Centre de Données astronomiques

de Strasbourg ———

Catalogues and Files available at the

Centre de Données Astronomiques

Version 1.5 January 19, 2008

Observatoire Astronomique

11, rue de l’Université 67000 – STRASBOURG, France

SPAN address: SIMBAD::QUESTION Internet address: question@simbad.u-strasbg.fr This document lists all files and catalogues available at CDS. The general organisation of the Archive of Catalogues and Published Data is made of the following sections:

I Astrometric Catalogues
II Photometric Catalogues (excluding Radio)
III Spectroscopic Catalogues
IV Cross-Identification
V Combined Data
VI Miscellaneous Catalogues
VII Non-stellar Objects
VIII Radio and Far-IR Catalogues
IX High-Energy Catalogues
J Published data (in Journals)

The catalogues are numbered in a chronological sequence within sections I to VIII, while the J section is order by Journal, volume, first page, where Journal is one of the following publications:

A+A  Astron. Astrophys.
A+AS  Astron. Astrophys. Suppl. Ser.
AJ  Astronom. J.
AZh  Astron. Zh.
ApJ  Astrophys. J.
ApJS  Astrophys. J. Suppl. Ser.
MNRAS Mon. Not. Royal Astron. Soc.
PASJ  Publ. Astron. Soc. Jap.
PASP  Publ. Astron. Soc. Pacific
PAZh  Pis’ma Astron. Zh.

Within the catalogue sections, the status line listed in the left column below the catalogue number indicates the availability of the data (data centers where a copy of the catalogue should exist, and availability over the network), with the following conventions:

B Bejing, WDC-D for Astronomy, Astronomical Observatory (China)
I India, Inter-University Center for Astronomy & Astrophysics, Pune
L La Plata, Centro de Datos Astronomicos (CDA), Argentina
M Moscow, Institute for Astronomy, Academy of Sciences, Russia
N NSSDC/WDC-A, NASA, Goddard Space Flight Center, Greenbelt, Maryland, USA
T Tokyo, Astronomical Data Analysis Center, Osawa, Japan
f fully available via anonymous ftp at cdsarc.u-strasbg.fr (also in FITS, see Ochsenbein 1995)

and on the Web at URL http://cdsweb.u-strasbg.fr/Cats.htmlS

v fully available on VizieR service at http://vizier.u-strasbg.fr

Further notes about the organisation of the Archive of Catalogues and Published Data can be found elsewhere (Ochsenbein 1992, 1993, 1994); a full description of the standards adopted for the documentation of the catalogues and tables at CDS is also available as the file /pub/cats/doc.tex on the anonymous ftp service installed on the cdsarc.u-strasbg.fr (130.79.128.5) Internet node, where the catalogues flagged with f here are also accessible.

This document is also available either in LaTeX source or in DVI format on the cdsarc.u-strasbg.fr

anonymous ftp service as file resp. /pub/cats/cats.tex or /pub/cats/cats.dvi

References

Ochsenbein, F., 1992, Bull. Inf. CDS 41, 65

Ochsenbein, F., 1993, Bull. Inf. CDS 42, 75

Ochsenbein, F., 1994, Bull. Inf. CDS 44, 19

Ochsenbein, F., 1995, Bull. Inf. CDS 46, 29

Catalogues available at CDS  ( January 19, 2008)  1

I. Astrometric Data

Catalogue  References  Contents  Remarks

AGK2/3 18Mb 183264 stars

I/2/

B-I-T

[Data in Simbad]

Lacroute P., Valbousquet A. , Astron. Astrophys. Suppl. Ser. 16, 345 (1974)

Id.: AGK2, BD m_{pg}, Sp α, δ, σ_{α}, σ_{δ} (1950) (Epochs 1930, 1958) μ_{α}, μ_{δ} , σ(μ)

Zone covered: −2.5^{∘} to +90^{∘} The plates have been reduced using an overlap method by Lacroute P. and Valbousquet A.

CAPE ZONE CATALOGUE 1.3Mb 20843 stars

I/5/

f-v

[Identifiers in Simbad]

SPENCER J.H., JACKSON J.: 1936, “Proper motions of stars in the Zone Catalogue of 20843 stars, Equinox 1900”, Her Majesty’s Stationery Office, London Id.: CZC, CPD m_{pg}, Sp α, δ (1900) μ_{α}, μ_{δ} Zone covered: −40^{∘} to −52^{∘}

Catalogue of 20554 Faint Stars in the Astrographic zone −40^{∘} TO −52^{∘} for the Equinox of 1900.0 1.3Mb 20554 stars

I/6/

T-f-v

[Identifiers in Simbad]

Spencer J.H., Jackson J., Her Majesty’s Stationery Office, London (1939)

Id.: CFC, CPD m_{pg}, Sp α, δ (1900) μ_{α}, μ_{δ} Zone covered : −40^{∘} to −52^{∘}

STERN-KATALOG FÜR DIE ZONE VON −6^{∘} BIS −10^{∘} SÜDLICHER DEK-LINATION FÜR DAS ÄQUINOKTIUM 1890, ERSTE UND ZWEITE ABTEILUNG 646Kb 10251 records

I/12/

B-I-M-N-T

[Identifiers in Simbad]

HERZ N.: 1907, Abh. des Königl. Preuss. Akademie der Wissenschaften vom Jahre 1906, Berlin Id.: Running n^{∘}, BD m_{pg} α, δ (1890)

Zone covered : −6^{∘} to −10^{∘} The positions are derived from meridian circle observations

PROPER MOTIONS OF 1160 LATE-TYPE STARS 49Kb 1160 stars

I/14/

B-I-M-N-T-f-v OLSEN H.J.F.: 1970, Astron. Astrophys. Suppl. Ser. 2, 69 Id.: HR, BD μ_{α}, μ_{δ} RV Stars observed photoelectrically by Dickow et al. in catalog II/38/

Katalog von 3356 schwachen Sternen für das Äquinoktium 1950 295Kb 3356 stars

I/16/

T-f-v Larink J., Hamburg-Bergedorf, Verlag der Sternwarte (1955)

Id.: Number, BD m_{pg}, Sp α, δ (1950) μ_{α}, μ_{δ} Zone covered: −5^{∘} to +90^{∘}

Catalogue  References  Contents  Remarks

KASAN CATALOGUE 100Kb 1144 records

I/17/

B-I-M-T NEFEDIAVA A.I.: 1963, Isvade 33, Isvetzia Astronomical Observatory Engelhardt Id.: Number, KSZ m_{pg} α, δ (1950)

Zone covered : +50^{∘} to +90^{∘}

MOSCOW CATALOGUE 133Kb 2154 records

I/18/

B-I-T GULYAEV A.P., KOROBOVA V.A., KHOMYAKOVA M.B.: 1970, Trudy Gaish 39, 21 Id.: Number, KSZ m_{pg} α, δ (1950)

Only stars of the KSZ. Zone covered: +30^{∘} to +45^{∘}

MERIDIANÖRTER VON 1570 STERNEN 114Kb 1567 records

I/19/

B-I-T HEINTZ W.D.: 1962, Veröff. Sternwarte München, 5, n^{∘} 15 Id.: BD, FK4, ADS, GC α, δ (1950) m_{pg} Zone covered : −30^{∘} to +90^{∘}

RESULTS OF OBSERVATIONS MADE WITH THE OTTAWA RE-VERSIBLE MERIDIAN CIR-CLE 1954-1962 134Kb 1557 records

I/20/

B-I-T WOOLSEY E.G.: 1972, Publ. Dominion Obs. Ottawa 25, n^{∘} 9 Id.: Number, BD, GC m_{pg}, Sp α, δ (1950)

Zone covered : −20^{∘} to +90^{∘}

CARTE  DU  CIEL

CATALOGUE: OXFORD + TOULOUSE + BORDEAUX + PARIS ZONES 81Mb 1696942 stars

I/21A/

N-T-f LACROUTE P., VALBOUSQUET A.: 1974, Bull. Inform. CDS 6 VALBOUSQUET A.: 1977, Bull. Inform. CDS 13, 2 LACROUTE P.: 1981, Bull. Inform. CDS 21, 2 Id.: zone + plate number + star number, AGK2 α, δ (B1950) plate epoch, m_{pg} Zone covered : +5^{∘} to +31^{∘} Zones in I/21/ and I/22/ are rearranged to I/21A/ and I/22A/. See also I/154/ and I/155/

CARTE DU CIEL CATA-LOGUE: ALGER ZONE 19Mb 306432 records

I/22A/

M-N-T-f LACROUTE P., VALBOUSQUET A.: 1974, Bull. Inform. CDS 6 VALBOUSQUET A.: 1977, Bull. Inform. CDS 13, 2 LACROUTE P.: 1981, Bull. Inform. CDS 21, 2 Id.: zone + plate number + star number, AGK2 α, δ (B1950) X, Y coordinate values Plate epoch, m_{pg} — Note: X, Y are not equal to those published.

Zone covered : −02^{∘} to +04^{∘} Zones in I/21/ and I/22/ are rearranged to I/21A and I/22A. Some data are missing and incomplete.

See also I/154/ and I/155/.

CATALOGUE OF PROPER MOTIONS FOR 437 STARS 23Kb 437 records

I/23/

B-I-M-N-T OLSEN H.J.F.: 1970, Astron. Astrophys. Suppl. Ser. 1, 189

Id.: HR μ_{α}, μ_{δ} V_{r} remarks

Catalogue  References  Contents  Remarks

AGK2 A 877Kb 13747 records

I/24/

B-I-T KOPFF A., PETERS J.: 1943, Astronomisches Rechen-Institut Veröff. 16, n^{∘} 71

Id.: ; m_{pg} ; α, δ (epoch) ; μ_{α}, μ_{δ}

ALBANY CATALOG OF 20811 STARS EPOCH 1910 866Kb 20811 records

I/25/

B-I-T BOSS B., ROY A.J., VAR-NUM W.B.: 1932, Carnegie Institution of Washington Id.: BD m_{pg} α, δ (1910) number of observations

BESANÇON: 2ème CATA-LOGUE 102Kb 1386 records

I/27/

B-I-T MAITRE V.: 1955, Ann. Obs. Besançon, fasc. IV Id. : BD α, δ (1930) σ(α, δ) number of observations

BONN 10: KATALOG VON 10663 STERNE 600Kb 10663 records

I/28/

B-I-N-T KÜSTNER F.: 1908, Veröffentl. Königl. Sternwarte zu Bonn, 10

{m_{pg} \atop α, δ (1900)}

BORDEAUX 50: CATA-LOGUE DE 2024 ETOILES 105Kb 2024 records

I/29/

B-I-T C.N.R.S.(Paris): 1958,

Id.: BD ; m_{pg} ; α, δ (1950)

Zone covered : +11^{∘} to +18^{∘}

BRORFELDE CATALOGUE 227Kb 3710 records

I/30/

B-I-T OLSEN H.J.F.: 1973, Astron. Astrophys. Suppl. Ser. 9, 1 Id. : n^{∘}, BD, GC m_{pg}, Sp α, δ (1973) (original epoch)

Meridian Observations made in Brorfelde from 1967 to 1979

Bucharest Catalogue: Catalogue KSZ d’étoiles faibles pour 1950.0 211Kb 3940 stars

I/31/

T-f-v Bucharest, Bucharest (1972)

Id.: KSZ, DM

m_{pg}, Sp α, δ (1950) number of observations Zone covered: −11^{∘} to +11^{∘}

FIRST GREENWICH CATA-LOG OF STARS FOR 1925.0

256Kb 2643 records

I/33/

B-I-N-T LONDON: 1924, London Obs.

Id.: GC (Boss’s cat.)

m_{pg}, Sp α, δ (1925) μ_{α}, μ_{δ}, N(obs) annual precession secular variation

Catalogue  References  Contents  Remarks

Second Greenwich Catalog of stars for 1925.0 1.4Mb 10587 stars

I/34/

T-f-v Royal Observatory, Greenwich , London (1935)

Id.: No, BD m_{v}, Sp α, δ (1925), Epoch μ_{α}, μ_{δ} N(obs) precession secular variation Zone covered : +32^{∘} to +64^{∘}

SECOND NINE YEAR CATALOG OF STARS FOR 1900: ASTROGRAPHIC REFERENCE STARS 980Kb 10127 records

I/35/

B-I-N-T LONDON: 1909, London Id.: Number, BD m_{pg} α, δ (1900) μ_{α}, μ_{δ}

Catalogue de 964 étoiles 50Kb 964 stars

I/36/

B-I-N-T Fayet G. , Cercle meridien Ann. Bur. Long. Vol. XIII (1963)

Id.: BD m_{pg} α, δ (1910) number of obs.

Zone covered : +5^{∘} to +15^{∘}

PARIS 50 : CATALOGUE DE 3997 ETOILES 173Kb 3997 records

I/37/

B-I-N-T PARIS , (Unpublished)

Id.: BD m_{pg} α, δ (1950) number of obs.

Zone covered : +33^{∘} to +35^{∘}

TOKYO MITAKA CAT-ALOGUE OF EQUATORIAL STARS 1950.0 (TME) 599Kb 4135 records

I/38/

B-I-N-T TUZI K. : 1962, Tokyo Ann. (2) 8,1 Id.: KSZ, BD α, δ (1950) μ_{α}, μ_{δ} number of obs.

RESULTS OF OBSERVATIONS MADE WITH THE 6” T.C. 1949-1956 CATALOGUE OF 5965 STARS FOR 1950, CORREC-TIONS TO FK4, GC AND N30: CAT. WASH. 1950 478Kb 5965 records

I/39/

B-I-T ADAMS A.N., BESTUL S.M., SCOTT D.K.: 1964, Publ. U.S. Naval Obs., 2nd series, XIX, part 1 Id.: DM, FK4, GC, N30 m_{pg}, Sp α, δ (1950), N(obs) Precession, Epoch Difference in α, δ with FK4, GC, N30

WASHINGTON 20: Catalog of 9989 standard and intermediary stars 754Kb 10571 stars

I/40/

T-f-v Morgan H.R., Publ. USNO, 2nd series, vol. XIII, Washington (1933)

Id.: Seq, DM m_{pg}, Sp α, δ (1920), Epoch μ_{α}, μ_{δ}, N(obs)

Catalogue  References  Contents  Remarks

PFKSZ: Preliminary Catalogue of Fundamental Faints Stars 43Kb 587 stars

I/42/

T-f-v Zverev M.S., Polozhentsev D.D. , Publ. Pulkovo (2), 72 (#87 in ARI Veroëff 16)

Id.: PFSZ, BD m_{pg}, Sp α, δ (1950) σ(α, δ), epoch Precession μ_{α}, μ_{δ}

Second Catalog of Fundamental Stars Santiago-Pulkovo (SFP-2) 42Kb 668 stars

I/44/

N-T-f-v Baturina G.D., Bedin V.S., Gnevysheva K.G., Zverev M.S., Naumova A.A., Plugina A.I., Polozhentsev D.D., Polozhentseva T.A., Stepanova E.A., Tr. Glav. Astron. Obs. Pulkovo, ser. 2 86, 4–158 Id.: FK4 m_{pg} α, δ (1950) from FK4 α − α_{F K4}, σ(α) Epoch, N(obs)

Catalogue de Zimmerman 286Kb 2957 stars

I/45/

T-f-v Zimmerman N.W. , Publ. Pulkovo (2), 61 ( #85 in A.R.I. Veröff. 16 )

Id.: Number, GC, FK4 m_{v}, Sp α, δ (1950), Epoch μ_{α}, μ_{δ}

CATALOGUE BABELS-BERG 4 (PRAGER) 734Kb 8803 stars

I/46/

B-I-T PRAGER R.: , Veröff. Berlin-Babelsberg 4 (24 in A.R.I. Veröff.)

Id.: Number, BD α, δ (1920), Epoch Precession, secular variation μ_{α}, μ_{δ}, m

CATALOGUE OF MELCHIOR- DEJAIFFE 33Kb 440 records

I/47/

B-I-T MELCHIOR P., DEJAIFFE R.: 1973, Obs. Royal Belgique Commun. (B) 74 Id.: Number, BD, GC α, δ (1950) (GC) μ_{α}, μ_{δ}

CATALOGUE MERIDIEN DE STRASBOURG 1972 200Kb 2832 records

I/49/

B-I-N-T STRASBOURG , (Unpublished)

Id.: Zverev, FK4, GC α, δ (Epoch) σ(α, δ) N(obs)

3ème CATALOGUE DE TOULOUSE 492Kb 10070 records

I/55/

B-I-N-T PALOQUE E.: 1937, Toulouse Annals, 13 Id.: Number, BD α, δ (1950) N(obs)

Zone covered : +3^{∘} to +12^{∘}

Catalogue  de  14263

étoiles comprises entre +16^{∘} et +24^{∘} 559Kb 14263 stars

I/57/

B-I-N-T-f-v , Hendaye (1915)

Id.: Number, BD m_{v} α, δ (1900)

Zone covered:

+16^{∘} to +24^{∘}

Catalogue  References  Contents  Remarks

Catalogue of proper motions of 12590 faint stars in the +25^{∘} to −20^{∘} declination zone 422Kb 12590 stars

I/59/

T-f-v Gorel L.F. , Trudy Glav. Astron. Obs. Pulkovo, 2nd serie LXXX, 15-154 Id.: Number, BD m_{pg} μ_{α}, μ_{δ} , σ(μ_{α}, μ_{δ})

AGK3 16Mb 183145 stars

I/61B/

N-T-f-v Heckmann O., Dieckvoss W., Hamburg-Bergedorf (1975)

Id.: Number, AGK3, BD m_{pg}, Sp α, δ (1950), Epoch μ_{1} = μα cos δ, μ_{δ} μ(AGK3) − μ(AGK2)

Zone covered: 90^{∘} to -2^{∘} ARI Heidelberg tape not absolutely identical with the printed AGK3

PERTH 70: A CATALOGUE OF POSITIONS OF 24900 STARS 3.7Mb 24978 stars

I/62C/

N-T-f-v HØG E., VON DER HEIDE J.: 1976, Abh. Hamburger Sternw., IX.

Id.: SRS, DM, FK4S α, δ (1950), Mean Epoch μ_{α}, μ_{δ} and σ σ(α), σ(δ), N(obs) m_{v}, Sp, π, V_{r} by exchange from NASA-Goddard Zones covered: −90^{∘} to +5^{∘}

A CATALOGUE OF 1649 STARS OBSERVED IN BOR-DEAUX WITH A TRACK-ING PHOTOELEC-TRIC MERIDIAN MICROM-ETER 131Kb 1649 records

I/63/

B-I-T MAZURIER J.M., MANGENOT G., RE-QUIEME Y.: 1977, Astron. Astrophys. Suppl. Ser. 27, 467 Id.: Number, BD m_{v}, Sp α, δ (1950.0) N(obs), Epoch

Catalogue d’Abbadia de +5^{∘}15′ à −3^{∘}15′ 546Kb 13532 stars

I/65/

B-I-T-f-v , Abbadia Obs. (1914)

Id.: Number, BD α, δ (1900), Epoch N(obs)

Zone covered: +5^{∘}15′ to −3^{∘}15′

SECOND CATALOGUE DE L’OBSERVATOIRE DE BOR-DEAUX 1900 DE +09^{∘} à +17^{∘} 351Kb 7038 records

I/66/

B-I-T BORDEAUX: 1900, Bordeaux Obs.

Id.: Number, BD

α, δ, epoch m_{V} , N(obs)

Zone covered: +9^{∘} to +17^{∘}

CATALOGUE DE 10656 ETOILES DE REPERE DE LA CARTE DU CIEL, Zone +16^{∘} à +24^{∘} 479Kb 10656 records

I/67/

B-I-T PARIS: 1929, Paris Obs.

Id.: Number, BD α, δ (1900), Epoch N(obs)

Catalogue  References  Contents  Remarks

Positions and proper motions of 1981 stars in the vicinity of α Per 159Kb 1981 stars

I/68A/

B-I-N-T-f-v Fresneau A., Bull. Inform. CDS 18, 81 (1980)

Id.: Heckmann, BD m_{pg}, Sp α, δ, σ, Epoch (1950) μ_{α} + σ, μ_{δ} + σ Supersedes the microfiche I/68/

Supplemental BD Stars 32Kb 838 records

I/71/

T-f-v Warren W.H., Kress K., Astronomical Data Center Bull. 1, 19 (1980)

Id.: BD (number & letter) α, δ (1855) BD magnitude By exchange from NASA-Goddard

AGK3R MEAN POSITIONS AND PROPER MOTIONS 1.6Mb 20194 records

I/72/

B-I-L-M-T

[Data in Simbad]

CORBIN T.E.: 1978, IAU Coll. 48, 505 Id.: AGK3R, BD α, δ (1950) +σ μ_{α}, μ_{δ} +σ Epoch, Weight weight

HKE SYDNEY CATALOG 2.1Mb 20457 records

I/73/

B-I-M-N-T EICHHORN H., GOOGE W.D., LUKAC C.F., MUR-PHY J.K.: 1963, Sydney Obs.

Id.: CPD α, δ (1950) m N(image measurements)

Zone −48^{∘} to −54^{∘}

QUATRIEME CATALOGUE MERI-DIEN DE L’OBSERVATOIRE DE BESANCON 84Kb 953 stars

I/74/

B-I-M-N-T CREZE M., MAZODIER B., CLAIREMIDI J., COLIN J., CONSIDERE S., HILAIRE G., OBLAK E., PARISOT J.P., PUEL F., ANDREZ R., ATHANASE M., CHA-BOD D., GODET L., JE-ANDENANS P., MOUGINS B., VINCENT A., VIN-CENT M.: 1982, Astron. Astrophys. Suppl. Ser. 50, 147 Id.: HD, FK4 m_{v} N(obs) α, δ (1950) Mean epoch Observed dispersion Binarity indexes Mean difference with FK4 Differential meridian positions for Ftype stars Zone +15^{∘} to +45^{∘}

SECOND CAPE CATALOGUE FOR 1950.0 417Kb 7180 stars

I/75/

B-I-M-N-T STOY R.H.: 1968, Annals of the Cape Obs. XXIII Id.: GC, DM m_{v}, Sp α, δ with σ Precession coeff. Epoch μ_{α}, μ_{δ} Differences with GC

Third Santiago-Pulkovo Fundamental Stars Catalogue (SPF3) 34Kb 671 stars

I/78/

N-T-f-v Loyola P.R., Shishkina V.N., SFP3, Publ. Dept. Astronomy Univ. of Chile 2, 159 (1974)

Id.: SFP3, FK4, N30, GC α, δ (1950.0), Epoch m_{v} N(obs), σ By exchange from NASA-Goddard

Catalogue  References  Contents  Remarks

Lowell Proper Motion Survey of Northern Hemisphere: the G-numbered stars 723Kb 8989 stars

I/79/

T-f-v Giclas H.L., Burnham R.Jr, Thomas N.G., Flagstaff, Lowell Obs. (1971)

Id.: G ; α, δ (1950) ; μ, θ ; m_{pg}, m_{v}, UBV

See I/112/ (was I/41/)

Catalog of 5268 standard stars, 1950.0, based on the normal system N30 482Kb 5545 stars

I/80/

N-T-f-v MORGAN H.R.: 1952, Astron. Papers Amer. Ephemeris 13, Part III Id.: Number, DM, GC m_{pg}, Sp α, δ (1950) μ_{α}, μ_{δ}, σ By exchange from NASA-Goddard I/13/ + notes

First Santiago-Pulkovo Fundamental Stars Catalogue (SPF1) 57Kb 1043 stars

I/82/

N-T-f-v Anguita C., Carrasco G., Loyola P., Bedin V.N., Naumova A.A., Polojentsev D.D., Polojentseva T.A., Tavastherna K.N., Zverev M.S., SFB1, Publ. Dept. Astron., Univ. of Chile, 2, 181 (1975)

Id.: FK4 α, δ (1950) α − α(F K4), σ(α) Epoch, N(obs)

By exchange from NASA-Goddard See also I/43/

Second Cape Photographic Catalogue 1950.0 (CPC2) I. Provisional Catalogue of positions of stars in the Cape zone −40^{∘} to −52^{∘} 4.1Mb 51017 stars

I/85/

T-f-v Nicholson W., Penston M.J., Murray C.A., de Vegt Ch., Mon. Not. R. Astron. Soc. 208, 911 (1984)

Id.: CPC2, SAO, SRS α, δ (1950), Epoch, σ m_{V} N(plates) μ_{α}, μ_{δ}, σ Difference SRS/CPC2 Provisional values in zones −40^{∘} to −52^{∘} See also I/265

Sydney Southern Star Catalogue 1.6Mb 26926 stars

I/86A/

T-f-v King D.S., Lomb N.R., J. Proc. R. Soc. N.S.W. 116, 53; Sydney Obs. Paper 96 (1983)

Id.: m_{pg} α, δ (1950), epoch N(plates) μ_{α}, μ_{δ} The Sydney measured positions (115740 records) is available on special request

Luyten Half-Second Catalogue (LHS) 364Kb 4471 stars

I/87B/

T-f-v

[Identifiers in Simbad]

Luyten W.J., Minneapolis, Univ. Minnesota (1976)

Id.: LHS, other α, δ (2000, 1950) m_{R}, m_{pg}, Sp μ, θ New version of I/54A Superseded by I/279

Catalogue  References  Contents  Remarks

SANTIAGO 67 CATALOGUE

653Kb 7610 records

I/89A/

I-N-T CARRASCO G., LOYOLA P.: 1981, Univ. de Chile, Dep. de Astronomia, IV.

Id.: SRS/BS/FK4, DM m_{v}, Sp α, δ (1950), Epoch, N(obs)

σ, μ, remarks 9 corrections to the printed version by exchange from NASA-Goddard

Accurate Positions of 502 stars of Pleiades 36Kb 502 stars

I/90/

I-M-N-T-f-v Eichhorn H., Googe W.D., Lukac C.E., Murphy J.K., Mem. Roy. Astron. Soc. 73, 125 (1970)

Id.: Hertzprung m_{v}, color index α, δ, σ μ_{α}, μ_{δ} Difference in μ with Hertzprung’s (1974) Difference in α, δ, with Koenig’s (1924)

By exchange from NASA-Goddard

Survey of the Astrographic Catalogue from 1^{∘} to 31^{∘} of Northern Declination 26Mb 997311 stars

I/96/

T-f-v Fresneau A., Astron. J. 88, 1378 (1983); Proc. Sympos. Strasbourg, Sept. 1983, Statistical Methods in Astronomy, ESA SP-201, 17 Id.: AC α, δ (1950), mean epoch m_{pg} By exchange from NASA-Goddard Based on data from I/21/ and I/22/

PERTH 75: A CATALOGUE OF POSITIONS OF 2589 FK4 AND FK4S STARS 205Kb 2589 records

I/97/

B-I-M-N-T-f-v NIKOLOFF I., HOG E.: 1982, Bickley W. Australia, Perth Observatory Id.: FK4 m_{v}, Sp α, δ (1950) μ_{α}, μ_{δ} N(obs), epoch By exchange from NASA-Goddard

New Luyten Catalogue of stars with proper motions larger than 0.2arcsec/yr (NLTT), and First Supplement 4.0Mb 58845 stars

I/98A/

M-N-T-f-v Luyten W.J., Minneapolis, University of Minnesota (1979) Luyten W.J., Hughes H.S., Proper Motion Survey with the Forty-Eight Inch Schmidt Telescope. LV (1980)

Id.: NLTT α, δ (1950) m_{pg}, m_{R} Sp or color class μ, θ addenda & supplement flags See the Revised NLTT by Salim & Gould J/ApJ/582/1011

Catalogue  References  Contents  Remarks

Brorfelde Meridian Catalogues: Brorfelde photographic meridian catalogues and Brorfelde photoelectric meridian catalogues 1.4Mb 13777 positions

I/99/

B-I-M-N-T-f-v Lautsen S., Publ. Mindre Medd. Kbh. Obs. (1968) Olsen H.J. Fogh, Astron. Astrophys. Suppl. Ser. 9, 1(1973) Olsen H.J. Fogh, Helmer L., Astron. Astrophys. Suppl. Ser. 30, 349 (1977) Helmer L., Olsen H.J. Fogh , Astron. Astrophys. Suppl. Ser. 49, 13 (1982) Helmer L., Fabricius C., Einicke O.H., Thoburn C., Astron. Astrophys. Suppl. Ser. 53, 223 (1983) Helmer L., Fabricius C., Einicke O.H., Thoburn C., Morrison L.V., Astron. Astrophys. Suppl. Ser. 55, 87 (1984)

Photographic catalogue Id.: BD, AGK3, FK4, NPZT, GC, ADS, BDS m_{v}, Sp α, δ, N(obs) Epoch, Parallax factor . Photoelectric catalogue Id.: DM, AGK3, SAOC, HD, GC, FK4/FK4Suppl, Gliese, ADS m_{v}, Sp α, δ, N(obs) Epoch

Washington 6-inch transit circle Catalogues (210, 025, 150, 250, 350, 450, 550) 2.7Mb 35175 stars

I/100A/

N-T-f-v Hammond J.C. , Publ. U.S. Naval Obs. (2) 11, 665 (1927) (W210) Watts C.B., Adams A.N., PUSNO2 16, Part I, 1 (1949) (W025); 16, Part I, 255 (1949) (W150) Watts C.B., Scott F.P., Adams A.N. , PUSNO2 16, Part III, 447 (1952) (W250) Adams A.N., Bestul S.M., Scott D.K., PUSNO2 19, Part I, 111 (1964) (W350); 19, Part II, 295 (1968) (W450) Hughes J.A., Scott D.K., PUSNO2 23, Part III, 165 (1982) (W450)

Id.: BD / CD ; α, δ (1950) ; μ_{α}, μ_{δ} ; N_{obs}

by exchange from NASA-Goddard Groups previous catalogs I/100 to I/106

Cape Photographic Durchmusterung (CPD: -18 to -89) 13Mb 454877 stars

I/108/

B-I-M-N-T-f-v Gill D., Kapteyn J.C., Annals of the Cape Obs., 3 – 5 (1895-1900)

Id.: CPD ; m_{pg} ; α, δ (1875)

By exchange from NASA-Goddard

Catalogue of double star occultation observations 54Kb 224 occultations

I/110/

N-T-f-v Evans D.S., Proc. IAU Coll. 62 Lowell Obs. Bull. 167, 73 (1983)

Id.: SAO, DM, name

m, Sp, ref.

Separation, σ(sep) PA, slope, σ(slope) B2 − B1, R2 − R1 Keypunched at the CDS Previously number VI/21/

Catalogue  References  Contents  Remarks

Southern Hemisphere Lowell Proper Motion Catalog 288Kb 2758 stars

I/112/

T-f-v GICLAS H.L., BURNHAM R. Jr., THOMAS N.G.: 1978, Lowell Obs. Bull. 164 Id.: Giclas α, δ (1950) μ, θ m_{pg}, color references

See I/79/

General Catalogue of 33342 stars (Version 1985) 5.0Mb 33342 stars

I/113A/

N-T-f-v Boss B. , Carnegie Inst. of Wash. Publ. 468 (1937)

Id.: GC, DM, SAO, HD m_{v}, Sp α, δ (1950); ℓ, b μ_{α}, μ_{δ}, σ annual and secular variations, 3rd term by exchange from NASA-Goddard supersedes I/113/

Cordoba Durchmusterung (CD: −22^{∘} to −89^{∘}) 18Mb 613959 stars

I/114/

B-I-M-N-T-f-v Thome J.M., Resultados del Observatorio National Argentino 16, 17, 18, parts I & II (1892-1932)

Id.: CD ; m_{v} ; α, δ (1875)

Contains I/88/ By exchange from NASA-Goddard

Cape Photographic Catalogue

4.9Mb 68467 stars

I/116/

T-f-v

[Identifiers in Simbad]

Stoy R.H., Jackson J. , Annals of the Cape Observatory 17 – 22 (1954–1968)

Id.: CPC, DM m_{pg}, m_{v}, Sp α, δ, σ(α, δ) (1950) μ_{α}, μ_{δ} Was I/4/ Zones covered: −30^{∘} to −40^{∘} −52^{∘} to −90^{∘}

ELEMENTS OF MINOR PLANETS FOR 1986 511Kb 3195 records

I/118/

B-M-N-T-f BATRAKOV Yu.V.: 1985, Izdatel’stvo Nauka, Leningrad Id.: Number, Name Epoch, mean anomaly, Argument of perihelion Node, Inclination, eccentricity, mean daily motion N(obs) σ, references by exchange from the Moscow Data Center see also I/124/ and I/127/

Südlicher Durchmusterung (SD: −1^{∘} to −23^{∘}) 4.4Mb 134833 stars

I/119/

N-T-f-v Schönfeld E., Astron. Beob. 8, part IV (1886)

Id.: SD (BD) ; m_{v} ; α, δ (1855)

By exchange from NASA-Goddard

Common Proper Motion Stars in the AGK3 116Kb 439 systems

I/121/

B-I-M-N-T-f-v Halbwachs J.L., Astron. Astrophys. Suppl. Ser. 66, 131 (1986)

Id.: AGK2, HD/DM, ADS/IDS, LDS, cluster m, Sp α, δ (1950) μ, θ, ρ

Catalogue  References  Contents  Remarks

Bonner Durchmusterung (BD: +89^{∘} TO −01^{∘}) 10.0Mb 325037 stars

I/122/

T-f-v Argelander F.W.A., Bonner Sternverzeichniss, 1. bis 3. Sektion, Astronomische Beobachtungen auf der Sternwarte der Kgl. Rhein. Friedrich-Wilhelms- Universität zu Bonn, 3–5 (1859-1862)

Id.: BD ; m_{v}(BD) ; α, δ (1855)

By exchange from NASA-Goddard supersedes I/83 and I/84

Lunar Occultations of IRAS Point Sources: 1986-1990 810Kb 14148 records

I/123/

B-I-M-N Simon M., Cassar L., Chen W.P., Astrophys. J. Suppl. Ser. 62, 673 (1986)

Id.: IRAS starting and ending years of the interval during which geocentric events occur for the object F (Jy), flags By exchange from NASA-Goddard see also I/142/

Elements of Minor Planets for 1987 563Kb 3330 records

I/124/

B-I-N-T Batrakov Yu.V. , Izdatel’stvo Nauka, Leningrad Id.: Number, Name Epoch, mean anomaly, Argument of perihelion Node Inclination, eccentricity, mean daily motion H, G (photometry) N_{obs}, N_{opp} References by exchange from the Moscow Data Center see also I/127/

COMBINED LICK-VOYAGER REFERENCE STAR CATALOGUE 2.0Mb 25136 stars

I/125A/

N-T-f-v

DUNHAM  D.W.  :

1986, Id.: Lick/DM, SAO, AGK3

α, δ (1950), Epoch, σ μ_{α}, μ_{δ}, σ m_{V} , m_{pg} π by exchange from NASA-Goddard

Elements of Minor Planets for 1988 810Kb 3516 records

I/127/

B-I-T BATRAKOV Yu.V., Izdatel’stvo Nauka, Leningrad Id.: Number, Name Epoch, mean anomaly, Argument of perihelion Node Inclination, eccentricity, mean daily motion H, G (photometry) N_{obs}, N_{opp} References by exchange from the Moscow Data Center see also I/124/

PHOTOGRAPHIC OBSER-VATIONS OF VISUAL DOU-BLE STARS 37Kb 272 observations

I/128/

N-T-f-v VAN ALBADA-VAN DIEN E., PANJAITAN E.: 1987, Astron. Astrophys. Suppl. Ser. 68, 117 Id.: ADS/IDS, name α, δ (1900, 2000), epoch m, Sp ρ, θ N(plates), notes

CATALOGUE OF 6125 STARS NEAR THE SOUTH GALACTIC POLE 784Kb 6181 stars

I/129/

B-I-M-N-T-f-v MURRAY C.A., ARGYLE R.W., CORBEN P.M. : 1986, Mon. Not. R. Astron. Soc. 223, 629 Id.: Number, RGO B, V, B − V α, δ, μ, π quality code, standard errors mean epoch X, Y See also J/AJ/103/514

Catalogue  References  Contents  Remarks

THE  LDS  CAT-

ALOGUE: DOUBLE STARS WITH COMMON PROPER MOTION 701Kb 6121 systems

I/130/

B-I-M-N-T-f-v LUYTEN W.J.: 1940,, Publ. Astr. Obs. Univ. Minnesota III, part 3, 35 LUYTEN W.J.: 1969–1987, Proper motion survey with the 48-inch Schmidt Telescope, XXI, XXV, XIX, XL, L, LXIV, LV, LXXI, Univ. Minnes.

Id.: LDS, LP, L, DM

α, δ (1950)

m_{pg}, m_{R}, color class

μ, θ

Position of component Remarks, Reference keypunched at the C.D.S.

Smithsonian Astrophysical Observatory (SAO) Star Catalog (edition ADC 1989) 51Mb 258997 stars

I/131A/

N-T-f-v

SAO Staff,  Publ.  of

the Smithsonian Institution of Washington, D.C. 4652 (1966)

Id.: SAO, DM, HD, GC α, δ (B1950, J2000), σ m_{pg}, m_{v}, Sp μ_{α}, μ_{δ} sources, remarks, By exchange from NASA-Goddard Errors discovered up to May 1989 were corrected

Catalogue of Trapezium Multiple Systems 57Kb 412 systems

I/134/

T-f-v SALUKVADZE G.N.: 1978– 1987 , Abastumans. Astrofiz. Obs. Bull. 49, 39 Astrofizika 14, 57 (1978) Bull. Inf. C.D.S. 33, 127 (1987)

Id.: IDS, ADS, DM

α, δ (1900)

ρ, θ m, Sp By exchange from the Moscow Data Center

CATALOGUE OF MULTI-PLE AND DOUBLE STARS OF T-ASSOCIATIONS 39Kb 525 records

I/135/

B-I-M-N-T SALUKVADZE G.N. : 1980, Astrofizika 16, 505; 16, 687 Id.: association, star α, δ (1900) m, Sp, colour indices ρ, θ By exchange from the Moscow Data Center

CATALOGUE OF TRAPEZ-IUM TYPE MULTIPLE SYSTEMS IN T-ASSOCIATIONS 46Kb 593 records

I/136/

B-I-M-N-T SALUKVADZE G.N.: 1980, Astrofizika 16, 505; 16, 687 Id.: association, star α, δ (1900) m, Sp, colour indices ρ, θ By exchange from the Moscow Data Center

GENERAL CATALOGUE OF POSITIONS AND PROPER MOTIONS OF 4949 GEODE-TICAL STARS FROM +90 TO -90 (CGS) 379Kb 593 records

I/137/

B-I-M-N-T KHRUTSKAYA E.V.: 1984 , Dep Union Inst. Scient. and Tech. Inf. 4077-84, 4078-84 Id. : CGS/1/, BS/FK5, BD, GC α, δ (2000), σ μ, m, Sp By exchange from the Moscow Data Center

The Southern Reference Star Catalogue of 20488 Positions 3.4Mb 20488 stars

I/138/

B-N-T-f-v Smith C.A., et al., in IAU Symposium 141, Inertial Coordinate System on the Sky, ed. J.H. Lieske and V.K. Abalakin, 457 (1988)

Id.: SRS, DM

m, Sp

α, δ (B1950), N(obs)

mean Epochs μ_{α}, μ_{δ} (B1950) α, δ, μ_{α}, μ_{δ} (J2000) by exchange from NASA-Goddard

Catalogue  References  Contents  Remarks

Yale Zone Catalogue 21Mb 221338 stars

I/141/

T-f-v Schlesinger F., Barney I., Trans. Astron. Obs. Yale Univ. 11–32 (1939-1983)

Id.: AGK1, DM m_{pg}, m_{v}, Sp α, δ, σ (1950) μ_{α}, μ_{δ}, σ

LUNAR OCCULTATIONS OF IRAS POINT SOURCES: 1991-2000 1.5Mb 27998 records

I/142/

B-I-M-N-T CASSAR L., CHEN W.P., SIMON M.: 1989 , Astrophys. J., Suppl. Ser. 69, 651 Id.: IRAS starting and ending years of the interval during which geocentric events occur for the object F (Jy), flags IRAS association By exchange from NASA-Goddard see also I/123/

FK4 and FK4 Supplement 778Kb 1535 stars

I/143/

T-f-v Fricke W., Kopff A., Veröff. Astron. Rechen-Inst., Heidelberg, 10, 11 Id.: FK4, GC, N30, FK4Sup α, ∂α/∂t, ∂^{2}α/∂t^{2}, σ(α) μ_{α}, ∂μ_{α}/∂t, σ(μ_{α}) δ, ∂δ/∂t, ∂^{2}δ/∂t^{2}, σ(δ) μ_{δ}, ∂μ_{δ}/∂t, σ(μ_{δ}) π By exchange from NASA-Goddard

First, Second and Third Herstmonceux catalogues of stars for 1950.0 4.5Mb 27493 stars

I/144/

B-I-M-N-T-f-v Tucker R.H., Buontempo M.E., Gibbs P., Swifte R.H.D., Royal Greenwitch Obs.Bull., 189 (1983)

Id.: AGK3R, PZT, Blaauw, GCRV, GCVS, Robertson, NPZT, GC, HD, SAO, AGK, ADS, IDS α, δ, epoch, N(obs) parallax factors, errors m, Sp, source source catalogue of mag. and sp. GC/HD/DM/SAO/AGK numbers

Catalog of Positions and Proper Motions (PPM-North) 47Mb 181731 stars

I/146/

B-I-M-N-T-f-v Röser S., Bastian U., Astron. Astrophys. Suppl. Ser. 74, 449 (1988) ident.: PPM, DM, AGK3, HD, SAO m_{v}, spectral type α, δ (1950/FK4 and 2000/FK5) p.m., epoch See also I/193/ and I/206/ By exchange from NASA-Goddard

FOURTH MELBOURNE GENERAL CATALOGUE OF 5157 STARS FROM OBSER-VATIONS 1894.0–1910.0 238Kb 5157 records

I/148/

B-I-M-N-T Melbourne Obs., Id.: Number α, δ (1900), Epoch m, N(obs)

−64^{∘} to −90^{∘}

Fifth Fundamental Catalogue (FK5). Part I. Basic fundamental stars 301Kb 1535 records

I/149A/

B-I-M-N-T-f-v Fricke W., Schwan H., Lederle T., Veröffentl. Astron. Rechen-Institut Heidelberg 32 (1988)

Id: FK5, GC, HD, DM, name m_{v}, Sp α, δ, μ (J2000) α, δ, μ (B1950) Mean epochs σ(α, δ, μ)

See also I/175/ By exchange from NASA-Goddard

Catalogue  References  Contents  Remarks

THE CATALOGUE OF PROPER MOTIONS FOR HIGH LUMI-NOSITY STARS 43Kb 309 stars

I/150/

N-T-f-v VALITOVA A.M., KARI-MOVA D.K., PAVLOVSKAYA E.D.: 1990, Trudy Gos. Astr. Inst. Shternberg Id.: number, HD Sp, m_{V} α, δ (J2000, FK5 epoch) μ, σ(μ) Epoch of observation By exchange from the Moscow Data Center

PRECISE RADIO SOURCE POSITIONS DETERMINED BY MARK III VLBI: OBSER-VATIONS FROM 1979 TO 1988 AND A TIE TO THE FK5 27Kb 182 radio sources

I/151/

N-T-f-v MA C., SHAFFER D.B., DE VEGT C., JOHNSTON K.J., RUSSELL J.L.: 1990, Astron. J. 99, 1284 Id.: IAU, other α, δ (J2000); σ_{α}, σ_{δ} α − δ correlation Mean Julian epoch N(experiments), N(obs) Other tables of reference by exchange from NASA-Goddard

Lick Jupiter-Voyager Reference Star Catalogue 230Kb 4983 stars

I/152/

B-N-T-f-v Klemola A.R., Morabito L., Taraji H., Lick Observatory, Univ. California, Santa Cruz(1978); corrections by Owen Jr W.M., 1990 Id.: star+plate number, AGK3 α, δ (1950) μ_{α}, μ_{δ} m_{V} , m_{pg} by exchange from NASA-Goddard

Lick Saturn-Voyager Reference Star Catalogue 203Kb 4551 stars

I/153/

T-f-v Klemola A.R., Taraji H., Ocampo A., Lick Observatory, Univ. California, Santa Cruz (1979)

Id.: star & plate numbers, other α, δ (1950) μ_{α}, μ_{δ} m_{V} , m_{pg} by exchange from NASA-Goddard

Astrographic Catalogue, zones -02 to +31 degrees: Right Ascension and Declination files 114Mb 2026347 stars

I/154/

N-T-f-v Röser S.: 1990, Astronomisches Rechen-Institut, Heidelberg (1990)

Id.: zone + plate number + star number, AGK3 α, δ (B1950) plate epoch, m_{pg} Zone covered: −02^{∘} to +31^{∘}. Re-calculation of I/21 and I/22. See also I/155. by exchange from NASA-Goddard

ASTROGRAPHIC CAT-ALOGUE, ZONES -02 TO +31 DEGREES: X,Y RECT-ANGULAR COORDINATES FILES 120Mb 2026564 records

I/155/

N-T-f RÖSER S.: 1990, Astronomisches Rechen-Institut, Heidelberg Id.: zone + plate number + star number X and Y coordinates on the plate plate epoch, image diameter Zone covered: −02^{∘} to +31^{∘}. See also I/154. by exchange from NASA-Goddard

The U.S. Naval Observatory Zodiacal Zone Catalog 4.8Mb 44435 stars

I/157/

B-N-T-f-v Douglass G.G., Harrington R.S. , Astron. J. 100, 1712 (1990)

Id.: DM α, δ (J2000, B1950) μ_{α}, μ_{δ} (J2000, B1950) by exchange from NASA-Goddard

Catalogue  References  Contents  Remarks

AGK3R 1.5Mb 21499 stars

I/161/

B-N-T-f-v Smith C.A., U.S. Naval Observatory (1990)

Id.: AGK3R, BD α, δ (1950), σ, epoch m_{pg}, Sp by exchange from NASA-Goddard

US Naval Observatory Pleiades Catalog 32Kb 506 stars

I/163/

N-T-f-v van Flandern T.C. , Unpublished (1969)

Id.: König ; α, δ (1950) ; P tv ; μ_{α}, μ_{δ}

by exchange from NASA-Goddard See also I/90/

THE TOKYO PHOTOELEC-TRIC MERIDIAN CIRCLE CATALOG 86 519Kb 3974 stars

I/167/

B-I-M-N-T-f-v YOSHIZAWA M., SUZIKI S. : 1989, Publ. Nat. Astron. Obs. Japan, 1, No 2, 127 Id.: BD, AGK3, other m_{v}, Sp α, δ, σ, mean Epoch

PERTH 83 1.6Mb 12331 stars

I/168/

B-M-N-T HARWOOD D.N.: 1990, Perth Observatory Id.: number, SAO m_{v}, Sp α, δ (B1950), N(obs) μ_{α}, μ_{δ}, residuals Mean epochs Contains more stars than the printed version The results were corrected for a systematic error

A CATALOG OF SELECTED COMPACT RADIO SOURCES FOR THE CON-STRUCTION OF AN EX-TRAGALAC-TIC RADIO/OPTICAL REF-ERENCE FRAME 20Kb 233 records

I/169/

B-N-T ARGUE A.N., DE VEGT C., ELSMORE B., FANSELOW J., HAR-RINGTON R., HEMEN-WAY P., JOHNSTON K.J., KÜHR H., KUMKOVA I., NIELL A.E., WALTER H., WITZEL A.: 1984, Astron. Astrophys. 130, 191 Id.: IAU system, other α, δ (J2000) + σ m_{V} , optical ident. code z, F (5GHz) Remarks by exchange from NASA-Goddard

Astrographic Catalogue Reference Stars 60Mb 320211 stars

I/171/

N-T-f-v Corbin T.E., Urban S.E., Warren W.H. Jr., NASA, NSSDC 91-10 (1991)

Id.: ACRS, DM, AGK3, CPC2 α, δ (B1950, J2000), σ μ_{α}, μ_{δ} (FK4, FK5), σ m_{pg}, Sp by exchange from NASA-Goddard

International Reference Stars Catalog (IRS) 6.8Mb 36027 stars

I/172/

N-T-f-v Corbin T.E., Warren W.H. Jr., NASA, NSSDC 91-11 (1991)

Id.: IRS, DM, AGK3 α, δ (B1950, J2000), σ μ_{α}, μ_{δ} (FK4, FK5), σ m_{pg}, Sp by exchange from NASA-Goddard Supersedes I/160

Catalogue  References  Contents  Remarks

Fifth Fundamental Catalogue (FK5). Part II: FK5 Extension 587Kb 3117 stars

I/175/

B-M-N-T-f-v Fricke W., Schwan H., Lederle T., Veröffentl. Astron. Rechen-Institut Heidelberg 32 , Verlag G. Braun, Karlsruhe (1988)

Id. : FK5, AKG3R, SRS, HD, DM, GC α + μ_{α}, δ + μ_{δ} (J2000 + B1950) mean Epochs; σ’s m_{V} , Sp, π, V_{r} Part I is I/149A/ By exchange from NASA-Goddard

The AGK3U: An Updated version of the AGK3 31Mb 170464 stars

I/176/

B-N-T-f-v Bucciarelli B., Daou D., Lattanzi M.G., Taff L.G. , Astron. J. 103, 1689 (1992)

Id.: AGK3 α, δ , σ (B1950) μ_{α}, μ_{δ} , σ Mean Epoch αδ(GSC), σ, Epoch

CATALOGUE OF PROPER MOTIONS IN NGC 752 118Kb 1777 stars

I/177/

B-N-T PLATAIS I.: 1992 , Bull. Inf. CDS 40, 5 Id.: Number, Heinemann, Rohlfs & Vanysek, Stock xy , m_{pg}, μ_{x}μ_{y} Membership probability

CATALOGUE OF PROPER MOTIONS IN NGC 7209 275Kb 4290 stars

I/178/

B-N-T PLATAIS I.: 1992 , Bull. Inf. CDS 40, 7 Id.: Number, Lavdovsky, Schewick, Maevers xy , m_{pg}, μ_{x}μ_{y} Membership probability

THE CHINESE GEODETI-CAL STARS CATALOGUE (CGSC) 526Kb 4065 stars

I/179/

B-N-T WORKING GROUP OF CGSC: 1992 , Bull. Inf. CDS 41, 5 Id.: FK5 / FK4, GC, DM α, δ (J2000), Epoch, σ μ_{α}, μ_{δ}, σ(μ_{α}), σ(μ_{δ}) m_{V} , Sp, π, V_{r} by exchange from Beijing

General Catalogue of stars with the system of photoelectric astrolabes (GCPA) 206Kb 1579 stars

I/180/

B-N-T-f-v Working Group of GCPA , Bull. Inf. CDS 41, 7 (1992)

Id.: Number, GC, FK4 α, δ (B1975), Mean Epoch

m_{V} , Sp Position Corrections, σ by exchange from Beijing

SECOND GENERAL CATA-LOGUE OF OF STARS OB-SERVED WITH THE PHOTOELECTRIC AS-TROLABES (GCPA2) 187Kb 2606 stars

I/181/

B-N-T LU LIZHI: 1992 , Bull. Inf. CDS 40, 9 Publ. Beijing Astron. Obs. 17, 1 Id.: FK5 / FK4 / GC α, δ (J2000), Mean Epoch m_{V} , Sp Position Corrections, σ by exchange from Beijing

Catalogue  References  Contents  Remarks

FOURTH PRELIMINARY CAT-ALOGUE OF STARS OB-SERVED WITH THE PHO-TOELECTRIC ASTROLABE (PACP4) 153Kb 1969 stars

I/182/

B-N-T-f-v LU LIZHI: 1992 , Bull. Inf. CDS 40, 11 Publ. Beijing Astron. Obs. 17, 59 Id.: Number, FK5 / FK4 / GC α, δ (J2000), Mean Epoch m_{V} , Sp Position Corrections, σ by exchange from Beijing

THE GIANT BRANCH OF THE FORNAX DWARF SPHEROIDAL GALAXY 39Kb 1287 stars

I/185/

B-N-T SAGAR R., HAWKINS M.R.S., CAN-NON R.D.: 1990 , Mon. Not. R. Astron. Soc. 242, 167 Id.: R-number, other x, y (1950) V, B − V

CATALOG OF POSITIONS OF 1007 STARS OBSERVED IN 1985 WITH TOKYO PHO-TOELECTRIC MERIDIAN CIRCLE (PMC85) 135Kb 1007 stars

I/186/

B-N-T-f-v YOSHIZAWA M., SUZUKI S., FUKAYA R.: 1987 , Ann. Tokyo Astron. Obs., Univ. Tokyo, 2nd Ser. 21, 399 Id.: Number, BD, AGK3, other α, δ (2000), σ, Mean Epoch m_{V} , σ(m_{V} ), Sp by exchange from NAO-Japan

CATALOG OF POSITIONS OF 5748 STARS OBSERVED IN 1987 WITH TOKYO PHO-TOELECTRIC MERIDIAN CIRCLE (PMC87) 743Kb 5748 stars

I/187/

B-N-T-f-v YOSHIZAWA M., SUZUKI S.: 1991 , Publ. National Astron. Obs. Japan, 2, 103 Id.: Number, BD, AGK3, other α, δ (J2000), σ Mean Epoch m_{V} , σ(m_{V} ), Sp by exchange from NAO-Japan

CATALOG OF POSITIONS OF 3800 STARS OBSERVED IN 1988 AND PLANETARY POSITIONS OBSERVED IN 1986 to 1988 WITH TOKYO PHOTOELECTRIC MERID-IAN CIRCLE (PMC88) 572Kb 3800 stars + 664 planet positions

I/188/

B-N-T-f-v YOSHIZAWA M., SUZUKI S., SOMA M.: 1992 , Publ. National Astron. Obs. Japan, 2, 475 Id.: Number, BD, AGK3, or Planet Name α, δ (J2000), σ Mean Epoch m_{V} , σ(m_{V} ), Sp by exchange from NAO-Japan

THE ASTROMETRIC STAN-DARD REGION NEAR THE NORTH GALACTIC POLE: A CATALOGUE OF POSI-TIONS AND PROPER MO-TIONS OF 3196 STARS 183Kb 3196 stars

I/192/

B-N-T PAKULYAK L.K.: 1992 , Bull. Inf. CDS 41, 21 Id.: Running number α, δ (1950), σ, Epoch m_{B}, μ_{α}, μ_{δ} Center 12h24 + 26^{∘}

by exchange from the Moscow Data Center

Catalogue  References  Contents  Remarks

PPM-South Star Catalogue 51Mb 197179 stars

I/193/

B-N-T-f-v Bastian U, Röser S., Yagudin L.I., Nesterov V.V., Polozhentsev D.D., Potter K.H., Wielen R., Yatskiv Y.S., Astronomisches Rechen-Institut, Heidelberg (1993)

Id: PPM, DM, AGK3, HD, SAO m_{v}, Sp α, δ (J2000 and B1950) μ_{α}, μ_{δ}, epoch Supersedes I/159A/

See I/146/ and I/208/

FOCAT-S: Pulkovo Photographic Catalogue of Southern Hemisphere 17Mb 201192 stars

I/195/

M-N-T-f-v Bystrov N.F., Polojentsev D.D., Potter H.I., Yagudin L.I., Zalles R.F., Bull. Inf. CDS 44, 3 (1994)

Id.: DM, SAO, PPM α, δ (J2000), σ, Epoch μ_{α}, μ_{δ} (J2000), σ N(obs) m, Sp

Hipparcos Input Catalogue, Version 2 54Mb 118209 stars

I/196/

I-N-T-f-v Turon C., Egret D., Gómez A., Grenon M., Jahreiß H., Réquième Y., Argue A.N., Bec-Borsenberger A., Dommanget J., Mennessier M.O., Arenou F., Chareton M., Crifo F., Mermilliod J.C., Morin D., Nicolet B., Nys O., Prévot L., Rousseau M., Perryman M.A.C , Bull. Inf. CDS 43, 5 (1993)

Id.: HIC, DM, HD, FK4-5, IRS, SAO, other α, δ (J2000, B1950) Epoch μ_{α}, μ_{δ}, σ, source H_{p} (Hipparcos Mag.) V, B − V + σ, source Sp, source π, V_{r} Variability parameters Duplicity parameters Remarks

Supersedes I/191/

The Tycho Input Catalogue (TIC) 262Mb 3154204 stars

I/197A/

I-N-T-f-v Egret D., Didelon P., McLean B.J., Russell J.L., Turon C., Astron. Astrophys. 258, 217 (1992)

Id.: TIC (GSC) ;

CSI, HIC, other α, δ (J2000, Ep. 1990), σ B, V , σ 27 flag qualifiers m_{V} ≤ 12.1 Preliminary selection flag was added

THE TOKYO PMC CATA-LOG 89: CATALOG OF PO-SITIONS OF 3466 STARS OBSERVED IN 1989 WITH TOKYO PHOTO-ELECTRIC MERIDIAN CIR-CLE (PMC89) 454Kb 3466 stars

I/198/

N-T-f-v YOSHIZAWA M., SUZUKI S.: 1993 , Publ. Nat. Astron. Obs. Japan 3, 45 Id.: Number, BD, AGK3a, other α, δ (J2000), O − C, σ Mean Epoch m_{V} , σ(m_{V} ), Sp by exchange from NAO-Japan

Lick Northern Proper Motion Program: NPM1 catalog 12Mb 148940 stars

I/199A/

N-T-f-v Klemola A.R., Hanson R.B., Jones B.F., Astron. J. 94, 501(1993)

Id.: NPM1, AGK3, SAO α, δ (B1950), Epoch Absolute proper motions B, B − V Discrepancy flags Supersedes I/11/ See I/200, I/214, I/283 by exchange from NASA-Goddard

Catalogue  References  Contents  Remarks

Lick Northern Proper Motion Program: NPM1 Reference Galaxies 2.4Mb 50517 galaxies

I/200/

N-T-f-v Klemola A.R., Hanson R.B., Jones B.F., Astron. J. 94, 501 (1993)

Id.: NPM1 α, δ (B1950 Epoch & Equinox) B (photographic) Discrepancy flags See I/199 by exchange from NASA-Goddard

The FK5 Extension on the FK4 System 552Kb 3117 stars

I/202/

N-T-f-v Lattanzi M.G., Taff L.G. , Astron. J. 105, 2353 (1993)

Id.: FK5 α, δ (B1950), Epoch μ_{α}, μ_{δ} by exchange from NASA-Goddard

Catalogue of Proper Motions, UBV-photometry and Spectral Classification in the region of NGC 7092 (M39) 928Kb 7931 stars

I/203/

N-T-f-v

[Data in Simbad]

Platais I. , Bull. Inf. CDS 44, 9 (1994)

Id.: Lavdovsky, Artiukhina, McNamara xy , α, δ (J2000) μ_{x}μ_{y} (relative) with σ α, δ (1950) UBV photometry Membership Pr.

THE BRIGHT STARS SUP-PLEMENT to the PPM and PPM SOUTH STAR CATA-LOGUES 51Kb 275 stars

I/206/

M-N-T-f-v BASTIAN U., RÖSER S.: 1993, Astronomisches Rechen-Institut, Heidelberg Id.: PPM m_{v} α, δ (J2000, Ep. 1950) μ_{α}, μ_{δ} Supersedes I/194 See I/146/ and I/193/

The 90000 stars Supplement to the PPM Star Catalogue 11Mb 89676 stars

I/208/

M-N-T-f-v Röser S., Bastian U., Kuzmin A., Astron. Astrophys. Suppl. Ser. 105, 301 (1994)

Id: PPM, DM, AGK3, HD, SAO m_{v}, Sp α, δ (J2000 and B1950) μ_{α}, μ_{δ}, epoch See also I/146 and I/193

Catalogue de 2700 étoiles doubles 1.0Mb 2714 double stars

I/209A/

N-T-f-v Couteau P., Observatoire de la Côte d’Azur, Département Augustin Fresnel, U.R.A. 1361, 2ème édition (Juin 1993)

Id.: COU, other α, δ (1900, 1950, 2000) m_{A}, m_{B} ; (ρ, θ) Remarks Supersedes I/173/ A few newer observations from I/209/

Improved mean positions and proper motions for the 995 FK4Sup stars not included in the FK5 Extension 207Kb 995 stars

I/210/

N-T-f-v

[Identifiers in Simbad]

Schwan H., Bastian U., Bien R., Jährling R., Jahreiß H., Röser S, Veröff. Astron. Rechen-Inst. Heidelberg, 34 (1993)

Id.: FK4, AG, SRS, HD, DM, GC α, δ; μ_{α}, μ_{δ} (J2000, B1950)

Mean epochs, σ(α) σ(δ), σ(μ) m_{V} , Sp, V_{r}, π

Catalogue  References  Contents  Remarks

CATALOGUE of PROPER MOTIONS in the region of NGC 3680. 323Kb 2711 stars

I/212/

N-T-f-v KOZHURINA-PLATAIS V., GIRARD T.M., PLATAIS I., VAN ALTENA W.F., IANNA P.A., CAN-NON R.D., Bull. Inf. CDS 46, 9 Id.: Anthony-Twarog, Eggen X, Y ; α, δ (J2000) Relative μ_{X}, μ_{Y} with σ BV (photographic) Membership probability

NPM1 Cross-Identifications and Appendices 3.8Mb 41858 entries

I/214/

f-v Hanson R.B., Klemola A.R., Jones B.F. , Bull. American Astron. Soc. 184, #27.07 (1994)

Id.: NPM1 , other Class See also I/199

TABLES OF MEMBERSHIP FOR 43 OPEN CLUSTERS (1994 Version) 1.7Mb 43 clusters

I/215/

M-N-T-f ZHAO J.L., TIAN K.P., Bull. Inf. CDS 47, 5 Id.: according to cluster X, Y ; μ_{X}, μ_{Y} by exchange from Beijing

Positions, Proper Motions, and Magnitudes in the area of the Open Cluster Tr 10 54Kb 979 stars

I/216A/

T-f-v STOCK J., Rev. Mex. Astron. Astrofiz. 9, 127 (1984)

Id.: Running number α, δ (1950, epoch 1982.4) V, m_{pg} from CIDA plates μ_{α}, μ_{δ}

The Magellanic Catalogue of Stars - MACS 12Mb 243561 stars

I/221/

M-N-T-f-v Tucholke H.-J, de Boer K.S. Seitter W.C., Astron. Astrophys. Suppl. Ser. 119, 91 (1996) The Messenger 81, 20 (1995)

Id.: MACS α, δ (J2000) Internal magnitude Position & Magnitude Flags

Stars of the southern hemisphere proposed for the Hipparcos mission 19Mb 96800 stars

I/230/

N-T-f-v Rousseau J.M., Perie J.P., Gachard M.T., Astron. Astrophys. Suppl. Ser. 116, 301 (1996)

Id.: PPM, HIC, SAO, HD, CPC, CPD, ADS/IDS, other α, δ (B1950, J2000) σ_{α}, σ_{δ}, Epoch, N_{obs} Mag, Sp Comments

The Washington Visual Double Star Catalog (WDS 1996.0) 6.3Mb 78100 binaries

I/237/

N-T-f-v Worley C.E., Douglass G.G., Astron. Astrophys. Suppl. Ser. 125, 523 (1997)

Id.: WDS, discoverer, DM α, δ (2000) Dates of first and last observations ρ, θ, N(obs) m_{A}, m_{B}, Sp, μ Notes & References Discoverers lists by exchange from NASA-Goddard Supersedes I/107

Catalogue  References  Contents  Remarks

Yale Trigonometric Parallaxes, Fourth Edition 1.3Mb 8112 stars

I/238A/

N-T-f-v van Altena W.F., Lee J.T., Hoffleit E.D., Yale University Observatory (1995)

Id.: GCTP, HR, HD, DM, name α, δ (1900) ; μ π+ σ, quality UBV, Sp by exchange from NASA-Goddard Supersedes I/174 and I/60 Errata by Erkki U. Vilkki

The Hipparcos and Tycho Catalogues 427Mb 1058332 stars

I/239/

T-f-v

ESA  1997,  ESA

SP-1200 Hipparcos Catalog (118218 stars): Id.: HIP, HD, DM α, δ μ_{α}, μ_{δ}, π(IRCS) Correlations αδμ_{α}μ_{δ}π Photometry B_{T} V_{T} Tycho Catalog (1058332 stars): Id.: TYC, PPM, HD, DM α, δ μ_{α}, μ_{δ}, π(IRCS) Correlations αδμ_{α}μ_{δ}π Photometry B_{T} V_{T} Variability & Multiplicity \underline{Annexes}: Double and Multiples Variability Solar System Result of HIPPAR-COS Mission Distributed on CD-ROMs by ESA

The N70E catalogue 64Kb 718 stars

I/241/

T-f-v Kolesnik Y.B., Mon. Not. R. Astron. Soc. 285, 1 (1997)

Id.: FK5 α, δ (2000) μ_{α}, μ_{δ}, + σ Epoch

{−30^{∘} < δ < \atop +30^{∘}}

The Pulkovo PVC96 catalogue: the first version of catalogue of FK5 stellar declinations observed in 1987–1994 with the Photographic vertical circle 48Kb 760 stars

I/242/

T-f-v Gontcharov A.G., Bagildinsky B.K., Kornilov E.V., Polojentsev D.D., Shkutov V.D., Izv. Glav. Astron. Obs., 213, 48 (1998)

Id.: FK5, HIC α, δ (J2000), σ Mean epoch from St Petersburg

OGLE (Optical Gravitational Lensing Experiment) General Catalog of Stars in BWC 2.1Mb 33196 stars

I/244A/

T-f-v Szymanski M, Udalski A., Kubiak M., Kaluzny J., Mateo M., Krzeminski W., Acta Astron. 46, 1 (1996)

Id.: OGLE-BWC α, δ (J2000) m_{I}, σ(m_{I}), V − I Remarks about microlensing candidates

Catalogue of Orbital elements and Photometric parameters of 7316 Minor Planets numbered by 25 November, 1996 1.4Mb 7316 planets

I/245/

T-f-v Batrakov Yu.V., Shor V.A., Institute of Theoretical Astronomy, St. Petersburg, Russia (1997)

Id.: Number, Name Epoch, mean anomaly, Argument of perihelion Node Inclination, eccentricity, mean daily motion H, G (photometry) N_{obs}, N_{opp} Element Sources & Residuals by exchange from St Petersburg Supersedes I/233/

Catalogue  References  Contents  Remarks

The ACT Reference Catalog 124Mb 988758 stars

I/246/

N-T-f-v Urban S.E., Corbin T.E., Wycoff G.L., U.S. Naval Observatory, Washington D.C.

Id.: AC2000, HD, HIP, TYC, DM α, δ (J2000), (μ_{α}, μ_{δ}) Tycho’s B_{T}, V_{T} Tycho’s flags by exchange from NASA-Goddard

The Tokyo PMC Catalog 90-93: Catalog of Positions of 6649 Stars Observed in 1990 through 1993 with Tokyo Photoelectric Meridian Circle 884Kb 6649 stars

I/248/

T-f-v Yoshizawa M., Suzuki S. , Publ. Nat. Astron. Obs. Japan 5, 1 (1997)

Id.: Number, DM, AGK3a, PZT, other α, δ (J2000), O − C, σ Mean Epoch m_{V} , σ(m_{V} ) by exchange from NAO-Japan

Catalog of O-B stars observed with Tokyo meridian circle 103Kb 1059 stars

I/249/

T-f-v Yasuda H., Hara H., Fukaya R., Ishii H., Ann. Tokyo Astron. Obs. 2nd Ser. bf 21, 107 (1986)

Id.: Seq, BD α, δ (1950) + σ, Epoch μ_{α}, μ_{δ} V, Sp

The Tycho Reference Catalogue 144Mb 990182 stars

I/250/

T-f-v Høg E., Kuzmin A., Bastian U., Fabricius C., Kuimov K., Lindegren L., Makarov V.V., Röser S., Astron. Astrophys. 335, L65 (1998)

Id.: TYC, HD, HIP, CCDM

α, δ (ICRS), μ_{α}, μ_{δ} σ(αδμ_{α}μ_{δ}) B_{T} , V_{T} ; σ(B_{T} V_{T}) π, flags

VLBI International Celestial Reference Frame (ICRF) 234Kb 608 sources

I/251/

T-f-v Ma C., Feissel M., IERS Tech. Note 23 (1997)

Id.: ICRF, IERS α, δ (J2000) Mean Epoch by exchange from NASA-Goddard

The PMM USNO-A2.0 Catalogue 120Kb 526280881 sources

I/252/

T-f-v Dave Monet, US Naval Observatory Flagstaff Station (1998)

Id.: USNO-A2.0 α, δ (2000) m_{R}, m_{B} (photographic) ACT(cat. I/246/) flag not accessible via FTP

The Guide Star Catalog Version 1.2 13Kb 25241730 records

I/254/

T-f-v Lasker B.M., Russel J.N., Jenkner H., The Association of Universities for Research in Astronomy, Inc. ([c] 1989, 1992)

Id.: GSC α, δ (J2000), σ m, σ(m) Multiplicity flag Plate The catalogue is accessible only for positional queries

Catalogue  References  Contents  Remarks

The Guide Star Catalog Version 14Kb 25241730 records

I/255/

T-f-v Lasker B.M., Russel J.N., Jenkner H., The Association of Universities for Research in Astronomy, Inc. ([c] 1989, 1992)

Id.: GSC α, δ (ICRS), σ m, σ(m) Multiplicity flag Plate The catalogue is accessible only for positional queries

Carlsberg Meridian Catalogs Number 1-11 105Mb 176591 stars

I/256/

f-v Copenhagen University Obs., Royal Greenwich Obs., Real Instituto y Observatorio de la Armada en San Fernando , Observations from 1984May to 1998May \underline{Program and FK5 stars}: Id: DM, SAO, GC, AGK3, HD, ADS, IDS α, δ (J2000), Ep., N_{obs} m_{v}, parallax factors \underline{Program stars only} : μ_{α}, μ_{δ}, σ_{μ}, Sp, DM, SAO \underline{Planets} : date, α, δ, m_{v} \underline{Variable stars} : name, date, m_{v} Supersedes I/147, I/170, I/189, I/205, I/213

Stars in Praesepe 116Kb 924 stars

I/257B/

T-f-v Wang J.J., Jiang P.F., Acta Astron. Sinica 32, 176 (1991) Wang J.J., Chen L., Zhao J.H., Jiang P.F. , Astron. Astrophys. Suppl. Ser. 113, 419 (1995)

Id.: Seq, Russel, other α, δ (J2000, Ep=1986.17) μ_{α}, μ_{δ} , σ V, B − V Membership Prob.

Pleiades positions and proper motions 42Kb 441 stars

I/258/

T-f-v Wang J., Chen L., Zhao J., Jiang P., Acta Astron. Sinica, 37, 68 (1996)

Id.: Seq, Hertzsprung α, δ (J2000), σ μ_{α}, μ_{δ}, σ

The Tycho-2 Catalogue 517Mb 2539913 stars

I/259/

T-f-v Høg E, Fabricius C, Makarov V V, Urban S, Corbin T., Wycoff G, Bastian U, Schwekendiek P, Wicenec A , Astron. Astrophys. 355, L27 (2000)

Id.: TYC, HIP, CCDM α, δ (ICRS J2000) σ_{α}, σ_{δ} (MeanEpoch) μ_{α}, μ_{δ} (ICRS) +σ ρ(α, δ) correlation Photometry B_{T} V_{T}

Catalogue of the Components of Double and Multiple Stars (CCDM) 5.4Mb 41255 components

I/260/

T-f-v Dommanget J., Nys O. , Astron. Astrophys. 363, 991 (2000)

Id.: CCDM , IDS , ADS, HD, DM, other α, δ (2000) ; μ_{α}, μ_{δ} ρ, θ, year of observation m, Sp

Supersedes I/211

The FON Astrographic Catalogue 131Mb 2008383 stars

I/261/

T-f-v Kislyuk V., Yatsenko A., Ivanov G., Pakuliak L., Sergeeva T., Main Astronomical Observatory of National Academy of Science of Ukraine (1999)

Id.: AC2000 α, δ (J2000); μ_{α}, μ_{d}elta BV R photometry Quality Flags

Catalogue  References  Contents  Remarks

ACR Catalog around the Celestial Equator 151Mb 1268732 stars

I/263/

T-f-v Stone R.C., Pier J.R., Monet D.G., Astron. J. 118, 2488 (1999)

Id.: ACR α, δ (ICRS) + σ R_{F} Epoch (mean J1996.0)

Sixth Catalogue of Fundamental Stars (FK6, parts I and III) 2.4Mb 4150 stars

I/264/

f-v Wielen R., Schwan H., Dettbarn C., Lenhardt H., Jahreiß H., Jährling R., Khalisi E., Veröff. Astron. Rechen-Inst. Heidelberg 35, 1 (1999); 37, 1 (2000)

Id. : FK5/6, HIP, other α + μ_{α}, δ + μ_{δ} (ICRS) mean Epochs; σ’s m_{V} , Sp, π, V_{r} Binarity Notes Extensive notes

Second Cape Photographic Catalogue (CPC2) 15Mb 274669 stars

I/265/

T-f-v Zacharias N., Zacharias M.I., De Vegt C., Astron. J. 117, 2895 (1999)

Id.: CPC2 α, δ (J2000) + σ, Epoch m_{pv} See the Preliminary Version I/85

Proper motions of fundamental stars (PMFS) 64Kb 1535 stars

I/266/

T-f-v Gontcharov G.A., Andronova A.A., Titov O.A., Kornilov E.V., Astron. Astrophys. 365, 222 (2001)

Id.: FK5, HIP μ_{α}, μ_{δ} (ICRS) Binarity status

The APM Catalogue 31Kb 166466987 sources

I/267/

f-v McMahon R. , Institute of Astronomy, Cambridge, CB3 OHA, UK Id.: APM α, δ (ICRS) Blue and Red Image classes, diameters Ellipticities, magnitudes

Catalog of Positions of IR Stellar Sources (CPIRSS) 9.4Mb 37700 sources

I/270/

T-f-v Hindsley R., Harrington R., Urban S. , Astron. J. 107, 280 (1994)

Id.: IRAS PSC, DM, WDS, HIP, TYC2 α, δ (J2000); μ_{α}, μ_{δ} IRAS fluxes (12, 25, 60, 100μm Astrometry flags Upgraded in 2001

M2000 : an astrometric catalog in the Bordeaux Carte du Ciel zone +11^{∘} < δ < +18^{∘} 234Mb 2286322 stars

I/272/

f-v Rapaport M., Le Campion J.-F., Soubiran C., Daigne G., Perie J.-P., Bosq F., Colin J., Desbats J.-M., Ducourant C., Mazurier J.- M., Montignac G., Ralite N., Requieme Y., Viateau B. , Astron. Astrophys. 376, 325 (2001)

Id.: M2000, or NLTT, G α, δ (J2000) + σ N_{obs}, Epoch V + σ

V_{M} <= 16.3

Catalogue  References  Contents  Remarks

Extragalactic Reference Link Catalog (ERLcat) 6.0Mb 89422 stars

I/273A/

f-v de Vegt C., Hindsley R., Zacharias N., Winter L., Astron. J. 121, 2815 (2001)

Id.: ERL α, δ (J2000) Epoch, N(plates)

Catalogue of the Components of Double and Multiple Stars (CCDM) 13Mb 49325 systems

I/274/

f-v Dommanget J., Nys O., Observations et Travaux 54, 5 (2002)

Id.: CCDM , IDS , ADS, HD, DM, other α, δ (2000) ; μ_{α}, μ_{δ} ρ, θ, year of observation m, Sp Id.: α, δ (J2000) σ Supersedes I/211 Was also numbered I/269

AC 2000.2: The Astrographic Catalogue on the Hipparcos System 425Mb 4621751 stars

I/275/

f-v Urban S.E., Corbin T.E., Wycoff G.L., Hoeg E., Fabricius C., Makarov,V.V., Astron. J. 115, 1212 (1998)

Id.: AC2000, ACRS, HIP, TYC α, δ (ICRS) + σ B_{T} , V_{T} Verification Flag

Supersedes I/247

The Tycho Double Star Catalogue 19Mb 65210 binaries

I/276/

f-v Fabricius C., Høg E., Makarov V.V., Mason B.D., Wycoff G.L., Urban S.E., Astron. Astrophys. 384, 180 (2002)

Id.: TDSC, TYC, WDS, HIP, HD α, δ (J2000), Epoch μ_{α}, μ_{δ} + σ B_{T} V_{T} ρθ

Yale/San Juan Southern Proper Motion Program: −20 to −40^{∘} zones including the South Galactic Pole Region 49Mb 321608 stars

I/277/

f-v Platais I., Girard T.M., Kozhurina-Platais V., van Altena W.F., Lopez C.E., Mendez R.A., Ma W.-Z., Yang T.-G., MacGillivray H.T., Yentis D.J. , Astron. J. 116, 2556 (1998)

Id.: SPM α, δ (J2000) + σ μ_{α}, μ_{δ} + σ Photographic BV Remarks

Catalogue de 1571 etoiles contenues dans l’amas globulaire Messier 3 (Catalogue of 1571 stars in the globular cluster M3) 164Kb 1571 stars

I/278/

f-v von Zeipel M.H., Ann. Obs. Paris 25, 1 (1908)

Id.: vZ α, δ (B1900), diff. plates m_{pg} Multiplicities Remarks

Revised LHS Catalogue 593Kb 4470 stars

I/279/

f-v Bakos G.A., Sahu K.C., Nemeth P. , Astrophys. J. Suppl. Ser. 141, 187 (2002)

Id.: LHS, Tycho-2 α, δ (J2000) μθ (recomputed) BV or m_{R}m_{pg} Contains 99% of the LHS (I/87)

Catalogue  References  Contents  Remarks

All-sky Compiled Catalogue of 2.5 million stars 415Mb 2501304 stars

I/280A/

f-v Kharchenko N.V., Kinematics and Physics of Celestial Bodies 17, 409 (2001)

Id.: ASCC-2.5, Tycho-2, HIP, HD, DM α, δ (J2000) + σ μ_{α}, μ_{δ} + σ BV (Johnson), scatter Sp Variability notes Errors in -1/+1^{∘} zone corrected in version A

Lick Northern Proper Motion Program: NPM2 Catalog 26Mb 232062 sources

I/283A/

f-v Hanson R.B., Klemola A.R., Jones B.F., Monet D.G. , Lick Observatory and US Naval Observatory, Flagstaff (May 2003); Astron. J. 128, 1430(2004)

Id.: NPM2, HIP, TYC, ACRS α, δ (J2000) μ_{α}, μ_{δ} Discrepancy flags

The USNO-B1.0 Catalog 34Kb 1045913669 sources

I/284/

f-v Monet D.G., Levine S.E., Casian B., et al. , Astron. J. 125, 984 (2003)

Id.: USNO-B1.0, TYC2 α, δ (ICRS) + σ μ_{α}μ_{δ} + σ m_{B}, m_{R}, m_{I} (DPOSS) Star/Galaxy class From POSS I + II

The GPM catalog 3.8Mb 52805 stars

I/285/

f-v Rybka S.P., Yatsenko A.I. , Main Astronomical Observatory of the National Academy of Sciences, Kiev, Ukraine, (2001)

Id.: KSZ region, TYC α, δ (J2000) μαμδ, Epoch BV (photographic) Statistics by region

The Second U.S. Naval Observatory CCD Astrograph Catalog (UCAC2) 2.7Mb 48330571 stars

I/289/

f-v Zacharias N., Urban S.E., Zacharias M.I., Wycoff G.L., Hall D.M., Germain M.E., Holdenried E.R., Winter L. , Astron. J. 127, 3043 (2004)

Id.: 2UCAC, 2MASS α, δ (ICRS) + σ μ_{α}, μ_{δ} + σ Epoch, N_{obs} m_{UCAC}, JHK (2MASS)

δ < 53^{∘}

Pul-3: A catalogue of positions and proper motions of 58483 stars on the Tycho-2 system 9.1Mb 58483 stars

I/290/

f-v Khrutskaya E.V.,Khovritchev M.Yu., Bronnikova N.M. , Izv. Glavnoj Astron. Obs. v Pulkove 216, 336; Astron. Astrophys. 418, 357 (2004)

Id.: Pul3, Tycho-2, HIP αδμ_{α}μ_{δ} (ICRS) + σ Mean Epoch B_{pg}R_{pg} List of plate centers

XZ Catalog of Zodiacal Stars (Version XZ80Q) 46Mb 244437 stars

I/291/

f-v Herald D., Canberra, Australia (August 2003)

Id.: XZ, DM, SAO, HIP, TYC, ZC, PPM, CCDM α, δ (J2000); μ_{α}μ_{δ} + σ Sp, BV Notes and flags Zodiacal strip |β| <= 6^{∘}40′ Supersedes I/201

Catalogue  References  Contents  Remarks

Pulkovo Visual Double Star Catalogue 135Kb 234 systems

I/292/

f-v Kiselev A.A., Kalinichenko O.A., Kiyaeva O.V., Shakht N.A., Romanenko L.G., Izmailov I.S., Bykov O.P., Maslennikov K.L., Bukhmastova Yu.L. , Leningrad, Nauka (2003)

Id.: WDS, ADS m_{A}m_{B}, Sp ρ, θ observations

NPM2 Cross-Identifications and Appendices 6.1Mb 46887 entries

I/293/

f-v Klemola A.R., Hanson R.B., Jones B.F., Monet D.G. , Astron. J. 128, 1430 (2004)

Id.: NPM2, Tycho-2, HIP Magnitudes, colors Class of object Notes and References

The UCAC2 Bright Star Supplement 91Mb 430000 stars

I/294/

f-v Urban S.E., Zacharias N., Wycoff G.L. , Astron. J. (unpublished)

Id.: 2UCAC, Tycho-2, 2MASS α, δ (J2000); μ_{α}, μ_{δ} BV IJHK Sources of data

Proper motions of 59766 stars with respect to galaxies 5.6Mb 59766 stars

I/295/

f-v Bobylev V.V., Bronnikova N.M., Schakht N.A., Pis’ma Astron. Zh.30, 519 (2004)

Id.: PUL2, Tycho, HIP α, δ, μ_{α}μ_{δ} (J2000) σ

Naval Observatory Merged Astrometric Dataset (NOMAD) 258Kb 1117612732 sources

I/297/

f-v Zacharias N., Monet D.G., Levine S.E., Urban S.E., Gaume R., Wycoff G.L. , San Diego AAS Meeting, January 2005 Id.: NOMAD1, USNO-B1, UCAC2, Tycho α, δ (J2000) + σ μ_{α}, μ_{δ} + σ BV RJHK Flags

A Catalog of Northern stars with proper motions larger than 0.15arcsec/yr 15Mb 61977 stars

I/298/

f-v Lepine S., Shara M.M. , Astron. J. 129, 1483 (2005)

Id.: LSPM, LHS, NLTT, TYC, UCAC2, 2MASS, USNO-B1 α, δ (J2000) μ_{α}, μ_{δ} (J2000) Astrometric Flag BV RIJHK

The PM2000 Bordeaux Proper Motion catalogue (+11^{∘} < δ < +18^{∘}) 679Mb 2670974 stars

I/300/

f-v Ducourant C., et al. , Astron. Astrophys. (in press)

Id.: PM2000, Tycho-2, 2MASS α, δ (J2000) +σ μ_{α}, μ_{δ} +σ V JHK N(obs)

Catalogue  References  Contents  Remarks

The XC1 catalogue of positions and proper motions of faint stars around the ICRF sources 89Mb 856421 stars

I/302/

f-v Fedorov P., Myznikov A. , Kinematics and physics of celestial bodies Suppl., 5, (2005)

Id.: Seq

α, δ (J2000) + σ

μ_{α}μ_{δ} + σ OEJF N (photographic)

The CdC2000 Bordeaux Carte du Ciel catalogue 25Mb 344781 stars

I/303/

f-v Rapaport M., Ducourant C., Le Campion J.F., Fresneau A., Argyle R.W., Soubiran C., Teixeira R., Camargo J.I.B., Colin J., Daigne G., Perie J.P., Requieme Y., Astron. Astrophys. 449, 435 (2006)

Id.: CDC2000, Tycho-2 α, δ (J2000), σ, Epoch m_{pg}

{Zone +11^{∘} < δ < \atop +18^{∘}}

Carlsberg Meridian Catalog Number 14 5.1Mb 95858475 stars (+8046291 rejected)

I/304/

f-v Copenhagen University Obs., Institute of Astronomy, Cambridge, Real Instituto y Observatorio de la Armada en San Fernando , Observations from March 1999 to October 2005 Id.: CMC14 α, δ (ICRS) + σ Ep(astrometry) r′JHK magnitudes Number of observations

−30 to +50^{∘} Supersedes I/282

The GSC 2.3 Catalog 32Kb 945592683 objects

I/305/

f-v , Space Telescope Science Institute (STScI) and Osservatorio Astronomico di Torino (2001)

Id.: GSC2.3

α, δ (ICRS), σ_{α}σ_{δ} m_{F} m_{J} m_{V} V m_{N} + σ Ellipse aeθ Class, Status From STScI Supersedes I/271 Proper motions not accessible

Catalogue of stars with highproper motions. V1. 54Mb 295525 stars

I/306/

f-v Ivanov G.A. , Main Astronomical Observatory (MAO), Kiev, Ukraine (2007)

Id.: FONAK1.0, AC2000, GSC α, δ (B1950, J2000) UBV RI, Sp V_{r} [Fe/H] References

Supersedes IV/28

Catalogue of 2601 stars. 159Kb 2601 stars

I/308/

f-v Gorel L., Publ. Pulkovo (2), 77, 23-73 (1969)

Id.: KSZ, BD α, δ (B1950) mvis

Catalogue of 5969 Southern Reference Stars and 726 Bright Stars. 404Kb 6695 stars

I/309/

f-v

Konin  V.V.,  Khrut-

skaya E.V., Publ. Pulkovo (2), 85, 4-148 (1983)

Id.: BS/SRS, BD α, δ (B1950) mvis

Catalogue  References  Contents  Remarks

The Catalogue of 5954 Stars in Declination Zone from -2deg to +1deg. 470Kb 5954 stars

I/310/

f-v Kortazzi I. , Catalog der Astronomische Gesellschaft, 1-121, (1900)

Id.: AGK1, BD α, δ (1875) Precession terms m_{v}

{Zone −2^{∘}10′ to \atop +1^{∘}10′}

II. Photometric Data

Catalogue  References  Contents  Remarks

Two Micron Sky Survey: a preliminary catalogue 1.0Mb 5612 sources

II/2B/

T-f-v

( microfiches ) [Data in Simbad]

Neugebauer G., Leighton R.B. , California Institute of Technology, NASA (1969)

Id.: IRC, BS, GC, DM, GCVS Sp α, δ (1950) KIV + σ magnitudes Zone covered: -30^{∘} to +80^{∘}

UBV Photometry of bright stars 387Kb 4777 stars

II/5A/

N-T-f-v Johnson H.L., Iriarte B., Mitchell R.I., Wisniewskj W.Z., Comm. Lunar & Planetary Lab. 4, 99, Table 9 (1966)

Id.: Geneva System UBV magnitudes These measures are included in UBV compilations See also II/122/

CELESCOPE CATALOGUE OF ULTRAVIOLET MAGNI-TUDES 1.0Mb 5761 records

II/6/

B-I-M-N-T

( microfiches ) [Data in Simbad]

DAVIS J.R., DEUTSCHMAN W.A., HARAMUNDANIS K.L.: 1973, SAO Special Report 350 Id.: Number, DM, NGC/IC, HD Sp α, δ (1950, 2000) U_{1}U_{2}U_{3}U_{4} magnitudes σ(U_{i}) Ambiguous identification flag Merged image flag Peculiarity The data are the same as those specified in the SAO Special Report 350, 31-32; CHECK-SUM was replaced by the running number of the record.

UBVRIJKLMNH Photoelectric Photometric Catalogue 776Kb 4494 stars

II/7A/

B-I-M-N-T-f-v

( microfiches ) [Data in Simbad]

Morel M., Magnenat P., Astron. Astrophys. Suppl. 34, 477 (1978)

Id.: Geneva System UBV RIJKLMNH N(obs) Remark Reference See also II/116/ CD-ROM ADC-1 Version

Catalogue of UVBGRI measurements: Published data and Weighted Means 191Kb 1297 stars

II/8/

N-T-f-v

Stebbins  J.,  Kron

G.E., Withford A.E., Strasbourg Information Bull. 2 (1971)

Id.: Geneva System UV BGRI magnitudes N(meas.) References Updated July 1975 See also II/75/

Catalogue of Photometric Measurements on the UBVr: Published data and Weighted mans 86Kb 366 stars

II/15/

B-I-M-N-T-f-v Magnenat P., Centre de Données Stellaires RISP, 6 (1973)

Id.: Geneva System Multiplicity (U − B)_{2}0 (B − V)_{2}0 (V − R)_{2}0 N(meas.) V Reference The system was defined by Sandage A. & Smith L.: Ap. J. 137, 1057 (1963)

Catalogue  References  Contents  Remarks

Catalogue des Mesures Photométriques dans le Système (102,65,62) de Eggen 91Kb 1585 stars

II/18/

B-I-M-N-T Magnenat P., Centre de Données Stellaires RISP 7, Avril 1974 Id.: Geneva System 102 65 – 62 102 – 65 N(meas.) References

Theoretical Colours for F and G dwarfs 11Kb 58 × 3 records

II/19/

B-I-M-N-T Bell R.A., Monthly Notices Roy. Astron. Soc 154, 343 (1971)

Id.: Number g, positions angle , vD Damping constant (M/H) ratio, T_{eff} UBV, uvby UV BRIJKL

The Spatial Distribution of Young Stars in Vela 53Kb 381 records

II/20/

B-I-M-N-T-f-v Denoyelle J. , Astron. Astrophys. Suppl. 27, 343 (1977)

Id.: Number, HD/CPD α, δ (1950) ; ℓ, b UBV, Sp Hα emission Q, E(B − V), A_{V} m − M, R Observations & Comments

Narrow-band Photometry of late-type stars 39Kb 629 stars

II/21A/

N-T-f Haggkvist L., Oja T. , Astron. Astrophys. Suppl. Ser. 1, 199 (1970)

Id.: HR, BD UBV data m_{4508}, g_{e}, c_{e} Remarks

HRA Photometry of Late-Type stars II. F and G dwarf stars South of the Equator 9.5Kb 172 records

II/22/

B-I-M-N-T Peat D.W., Mon. Not. Roy. Astron. Soc. 131, 467 (1966)

Id.: HD U_{c}BV α index

Hα Photometry of Late-Type Stars. I. F, G and K type stars North of the Equator 52Kb 594 stars

II/24/

B-I-M-N-T-f-v Peat D.W. , Mon. Not. Roy. Astron. Soc. 128, 435 (1964)

Id.: HD, MK (B − V) Remark Mg b, Na D, Hα ratios

Photometry of Orangered Ca-I Triplet in Late-Type Stars 24Kb 296 records

II/25/

B-I-M-N-T Peat D.W., Mon. Not. Roy. Astron. Soc. 128, 475 (1964)

Id.: HD (B − V) Mg b, Hα, CaI ratios

Catalogue  References  Contents  Remarks

CATALOGUE OF EARLY-TYPE STARS MEASURED IN A NARROW-BAND PHO-TOMETRIC SYSTEM 97Kb 1482 records

II/26A/

N-T-f MORGULEFF N., GER-BALDI M.: 1975 , Astron. Astrophys. Suppl. Ser. 19, 189 Id.: HD or BD C(j) − C(2) for j = 1, 3 ⋯ 8 σ(C_{j})

Catalogue of Photoelectric Photometric Measurements in the UcBV system 905Kb 7146 stars

II/27/

B-I-M-N-T Nicolet B., Astron. Astrophys. Suppl. Ser. 22, 239 (1975)

Id.: Geneva System U_{c}BV photometry N(obs), Reference Weighted Means It is kept up-to-date

uvbyβ Photometry for bright O- to G0-type stars south of declination +10^{∘} 297Kb 2828 records

II/33/

T-f-v

[Data in Simbad]

Gronbech B., Olsen E.H., Astron. Astrophys. Suppl. Ser. 25, 213 (1976, 1977) Astron. Astrophys. Suppl. Ser. 27, 443 Id.: HR V, uvbyβ σ, weights Remarks uvby photometry is the same as II/30/

The

data are included in II/57/

Catalogue of Individual UBV and uvbyβ Observations of Stars in the region of Orion OB-1 Association 134Kb 1976 records

II/36/

B-I-M-N-T-f-v Warren W.H., Hesser J.E., Astrophys. J. Suppl. Ser. 34, 115 (1977)

Id.: Number, Brun, P, HD, BD Obs. dates UBV Photometry uvbyβ Photometry

Catalogue of UBV Photometry of Globular Cluster Stars 2.0Mb 113 clusters

II/37/

N-T-f-v Philip A.G.D., Cullen M.F., White R.E., Dudley Obs. Report 11 (1976)

Id.: NGC/M

E(B − V) , F e/H , Sp V , B − V , U − B References

Catalogue Miniphote 425Kb

II/38/

I-M-N-T MagnenaT P., Bull. Inform. CDS 8, 20 (1975)

Id.: Geneva System

Photometric measurements in 13 systems N(meas.) References

Photoelectric Measures of Hydrogen-line absorption in Early-Type Stars 16Kb 177 records

II/40/

I-M-N-T Bappu M.K.V., Chandra S., Sanval N.B., Sinvhal S.D. , Mon. Not. R. Astron. Soc. 123, 521 (1962)

Id.: HD V_{0}, (U − B)_{0} Hγ intensities

Catalogue  References  Contents  Remarks

Photoelectric Measurements of the 4200Å CN band and the G band in G8-K5 Spectra 48Kb 712 records

II/41/

B-I-M-N-T Griffin R.F., Redmann R.O. , Mon. Not. R. Astron. Soc. 120, 287 (1960)

Id.: HD Sp CN ratio C band ratio Notes

K-line Photometry of A Stars

25Kb 146 stars

II/42/

B-I-M-N-T Henry R.C., Astrophys. J. Suppl. Ser. 18, 47 (1969)

Id.: HR, HD Stroemgren’s group Sp, K, δK Hβ, b − y, (b − y)_{0}, (c_{1}), (m_{1}) a (temperature parameter)

K-Line Photometry of Southern A stars: Summary of data on Southern A stars 49Kb 738 stars

II/43/

B-I-M-N-T-f-v Henry R.C., Hesser J.E., Astrophys. J. Suppl. Ser. 23, 421 (1971)

Id.: HR V, B − V K-line index +σ δk Stroemgren’s group b−y, Hβ, (m1), δm1

K-line Photometry of Stars in Population I Clusters 20Kb 112 entries

II/44/

B-I-M-N-T Hesser J.E., Henry R.C., Astrophys. J. Suppl. Ser. 23, 453 (1971)

Id.: cluster number V, B − V, m_{1}, c_{1} K–strength, σ(K), δK Stroemgren’s group [b − y], [m1], [c1], [u − b], Hβ A-temperature param.

Pleiades (32 stars) IC 2391 (24 stars) IC 2602 (32 stars) Hyades (24 stars)

Strong Cyanogen Stars 13Kb 185 records

II/45/

B-I-M-N-T Janes K.A., McClure R.D., Astrophys. J. 165, 561 (1971)

Id.: HD C(45 − 48), C(42 − 45), C(41 − 42) δ(U − B), δ(Cm) Sp, Distance The data are included in II/17/

A Photometric Investigation of Strong Cyanogen Stars 17Kb 183 records

II/46/

B-I-M-N-T McClure R.D., Astron. J. 75, 41 (1970)

Id.: HR, HD UBV, Sp C(45-48) C(42-45) C(41-42) Cm(41-42) C(42-48) unreddened Cm(41-42) unreddened The data are included in II/17/

Scanner abundance studies. I. An investigation of supermetallicity in late-type evolved stars 74Kb 229 stars

II/47/

B-I-M-N-T-f-v Spinrad H., Taylor B.J., Astrophys. J. 157, 1279 (1969)

Id.: HR 23 colours in range [3880 ⋯ 7400]Å Blocking fractions Remarks

Catalogue  References  Contents  Remarks

A Reference List for the UBV System 436Kb 13425 records

II/48/

B-I-M-N-T-f-v Nicolet B., Bull. Inform. CDS 11, 20 (1976)

Id.: Geneva System UBV (a weighted mean)

Photometric Standard Stars: magnitudes and colours of bright Southern stars & Stars in Equatorial zone 148Kb 900 records

II/50/

B-I-M-N-T Cousins A.W.J. , Royal Observ. Ann. 7 (1971)

Id.: HR UBV (weighted mean) (reduction as adopted in Lunar & Planetary Lab. Comm. 63

Catalog of 10μm Celestial Objects 136Kb 647 records

II/53/

B-I-M-N-T Hall R.T., Report for Space & Missile System Organisation: SAMSO-TR-74, 212 (1974)

Id.: Name, DM, SAO, HR, HD, GC, IRC α, δ (1950) ; ℓ, b μ_{α}, μ_{δ} V IKN magnitudes Fluxes References By exchange from NASA-Goddard

Gershberg Flare Star Catalogue 8.5Kb 53 stars

II/55/

N-T-f-v Shakhovskaya N.I., IAU Coll. 15 Veröff. Bamberg IX, 138 (1971)

Id.: Name α, δ (1950) Multiplicity Comments By exchange from NASA-Goddard

100μm Survey of the Galactic Plane 7.5Kb 72 records

II/56/

B-I-M-N-T Hoffmann W.F., Frederick C.L., Emery R.J., Astrophys. J. 170, L89 (1971)

Id.: Number α, δ (1950) ; ℓ, b F (100μm): peak, total Comments By exchange from NASA-Goddard

Catalogue of Ultraviolet Stellar Fluxes (TD1) 3.3Mb 31215 stars

II/59B/

M-N-T-f-v

( microfiches )

Thompson G.I., Nandy K. Jamar C., Monfils A. Houziaux L., Carnochan D.J., Wilson R., Science Research Council (1978)

Id.: HD, DM F (λ) centered 1565, 1965, 2365, 2740 Å See also III/39/ & III/53/. Microfiche produced by Univ. College London, (see Carnochan D.J.: 1979, Bull. Inform. CDS 17, 18)

Catalog of Extinction Data 2.4Mb 12547 records

II/62/

M-N-f-v Neckel Th., Klare G., Sarcander M., Bull. Inform. CDS 19, 61 (1980)

Id.: HD, DM, KS, SS, USN α, δ (1950) Sp, UBV, Hβ Distances, Extinctions References see also Astron. Astrophys. Suppl. Ser. 42, 251 (1980)

Catalogue  References  Contents  Remarks

A Compilation of Balmer lines photometric data 455Kb 11 photometries

II/63A/

T-f-v Mermilliod J.C., Mermilliod M., Bull. Inform. CDS 19, 65 (1980)

Id.: Geneva system

αβγδ photometric indices data from 11 photometric systems using Balmer lines N(meas.)

Five Color Photometry of Blue Stars in and between the Magellanic Clouds 86Kb 1128 records

II/65/

B-I-M-N-T Wamsteker W., Astron. Astrophys. Suppl. Ser. 43, 27 (1981)

Id.: Number α, δ (1975) V BUW L (Walraven) N(meas.)

A Catalogue of Observations in Hα 325Kb 2299 stars

II/66/

N-T-f-v Ducati J.R., Astron. Astrophys. Suppl. Ser. 45, 119 (1981)

Id.: HD m, Sp Hα Reference

Ultraviolet Spectrophotometric Catalogue from the OAO Satellite 344Kb 41 records

II/67/

B-I-M-N-T Code A.D., Meade M.R. , Astrophys. J. Suppl. Ser. 39, 195 (1979)

Asbolute energy distribution per unit wavelength interval between 1200Å and 3580Å

Dearborn Observatory Catalogue of faint red stars 1.4Mb 44076 stars

II/68/

T-f-v Lee O.J., Gore G.D., Baldwin T.J., Ann. Dearborn Obs. V, 1A-C (1947)

Id.: DO αδ (1900) m_{R} Sp(HD), Sp(DO)

By exchange from NASA-Goddard

Catalogue of Stars Suspected of Variability 1.3Mb 8904 records

II/69/

B-I-M-N-T Kukarkin B.V., Kholopov P.N. Efremov Y.N., Kurochkin N.E., Moscow (1951, 1965)

Id.: HV, ROSS m_{max}, m_{min} α, δ (1900) Variability type Confirmed variables were not eliminated.

By exchange from NASA-Goddard See also II/79/ Same items as II/12/

Catalogue of Dubious Variable Stars 155Kb 3137 records

II/70/

B-I-M-N-T Kukarkin B.V., Kholopov P.N., Efremov Yu.N., Kurochkin N.E. , Moscow (1951, 1965)

Id.: HD/DM m_{v}, Sp α, δ (1900) Var. type Confirmed variables were not eliminated.

By exchange from NASA-Goddard Same items as II/13/

Catalogue  References  Contents  Remarks

Alexander Narrow-Band Photometry 120Kb 1257 records

II/73/

B-I-M-N-T Alexander J.B. , Mon. Not. R. Astron. Soc. 194, 403 (1981)

Id.: star name α, δ (1900); ℓ, b m_{λ}, N(obs) Sp(derived from data)

Photometric Data for the Nearby Stars 102Kb 1570 stars

II/74/

N-T-f-v Hauck B., Mermilliod M. , Bull. Inform. CDS 21, 35 (1981)

Id.: Gliese, HD, DM m_{v} Photometric systems in which the star is measured Subset of II/167/

Homogeneous Catalogue of Red and Infrared Magnitudes in the Photoelectric Photometric system of Kron 961Kb 11989 records

II/75/

B-I-M-N-T-f-v Jasniewicz G., Astron. Astrophys. Suppl. Ser. 49, 99 (1982)

Id.: Geneva system R, R − I N(meas.), weights References

Catalogue of non-Solar X-ray Measurements 698Kb 1301 records

II/77/

B-I-M-N-T Arens J., Rothschild R. , NASA/GSFC X-661-75-230 (1975)

Id.: source name α, δ ; ℓ, b Energy range & FWHM Date of observation Exposure time Effective area of obs. Detector type Payload vehicle Reference & Comments By exchange from NASA-Goddard

uvbyβ Photometry of Multiple Stellar Systems 50Kb 398 records

II/78/

B-I-M-N-T-f-v Olsen E.H. , Astron. Astrophys. Suppl. Ser. 48, 165 (1982)

Id.: IDS, HD, DM V, b-y, m1, c1, H-beta n. of nights, weights and rms errors

Catalogue of Suspected Variable Stars 2.1Mb 14850 records

II/79/

B-I-L-M-N-T Kukarkin B.V. et al., Moscow (1981)

Id.: NSV, other α, δ (1950) m(max), m(min) UBRIK Sp, Variability type References By exchange from Moscow Data Center see also II/140/

A Catalogue of homogeneous photometry of bright stars in the DDO system 181Kb 2196 stars

II/80/

B-I-M-N-T-f-v McClure R.D., Forrester W.T., Publ. Dominion Astrophys. Obs. 15, 439 (1981)

Id.: HD, DM Magnitudes in 48 filters, standard errors By exchange from NASA-Goddard

Catalogue  References  Contents  Remarks

Catalogue of Galactic O-type Stars 74Kb 765 records

II/82/

B-I-M-N-T-f-v Garmany C.D., Conti P.S., Chiosi C., Astrophys. J. 263, 777 (1982)

Id.: HD, DM, other ℓ, b Sp, V, B − V M_{V} , M_{bol}, distance cluster or association membership By exchange from NASA-Goddard Distributed also in floppy-disk 5”1/4

OAO2 filter Photometry of 531 stars of diverse types 94Kb 531 stars

II/83/

T-f-v Code A.D., Holm A.V., Bottemiller R.L., Astrophys. J. Suppl. Ser. 43, 501 (1980)

Id.: HD, HR, name Sp, UBV m(λ) for λλ (Å): 1330, 1430, 1550, 1680, 1910, 2035, 2380, 2460, 2945, 2980, 3320 and 4250 References and remarks By exchange from NASA-Goddard

13-color Photometry of 1380 bright stars 178Kb 1380 records

II/84/

N-T-f-v Johnson H.L., Mitchell R.I. , Rev. Mex. Astron. Astrofis. 1, 299 (1975)

Id.: HR ; Sp 52, 33, 35, 37, 40, 45, 58, 63, 72, 80, 86, 99, 110 magnitudes By exchange from NASA-Goddard Same items as II/52/

Absolute Calibration of stellar spectrophotometry 20Kb 16 stars

II/85/

T-f-v Johnson H.L., Rev. Mex. Astron. Astrofis. 5, 25 (1980)

Id.: name, HR, HD F (λ) for 94 λ-values 4045 to 10220 Å By exchange from NASA-Goddard

Supplement to UV Bright Star Spectrophotometric Catalogue 638Kb 435 spectra

II/86/

M-T-f-v Macau-Hercot D., Jamar C., Monfils A., Thompson G.I., Houziaux L., Wilson R. , Supplement to the ultraviolet bright star spectrophotometric catalogue, ESA SR-28 (1987)

Id.: HD λ, F (λ) σF λ ∈ [1350, 2540] Å (step 10Å)

Supplement to III/39/

Photographic Magnitudes for stars at high Galactic latitude 810Kb 15684 records

II/87/

B-I-M-N-T-f-v Weistrop D., Bull. Inform. CDS 24, 91 (1983)

Id.: Number X, Y position V, B, U, r magnitudes

uvbyβ data for Local Interstellar Reddening Map 208Kb 3458 records

II/88/

B-I-M-N-T Perry C.L., Johnston L., Astrophys. J. Suppl. Ser. 50, 451 (1982)

Id.: HD ; ℓ, b ; V, uvbyβ

Catalogue  References  Contents  Remarks

Four-color uvby and Hβ Photometry of A5 to G0 Stars brighter than 8.3 1.5Mb 15481 records

II/90/

B-I-M-N-T Olsen E.H., Astron. Astrophys. Suppl. Ser. 54, 55 (1983)

Id.: HD or DM V, uvbyβ N(obs.), Weights, σ

The intrinsic Colours of stars in the U.V. 223Kb 882 stars

II/91/

T-f-v Carnochan D.J., Mon. Not. Roy. Astron. Soc. 201, 1139 (1982)

Id.: Number Sp, UBV, M_{V} (UV-V) color indexes λλ: w=330Å: 1565, 1965, 2365, 2740 w=110Å: 1455, 1565, 1675, 1855, 1965, 2075, 2255, 2365, 2475 By exchange from NASA-Goddard See also II/58

Revised S201 Catalog of Far-UV Objects 829Kb 6396 records

II/92/

B-I-M-N-T Page T., Carruthers G.R., Heckathorn H.M., NRL Report 8487 (1982)

Id.: SAO, HD X, Y ; α, δ Sp, m_{v} Exposure time ; Filter; 1050-1600 or 1250-1600 Å By exchange from NASA-Goddard

The Revised AFGL I.R. Sky Survey Catalog and Supplement 841Kb 6146 sources

II/94/

T-f-v

Price  S.D.,  Mur-

dock T.L., Air Force Geophysics Lab. AFGL-IR-83-0161 (1983)

Id.: AFGL, RAFGL αδ (1950) m(λ) at 4, 11, 20, 27 μm By exchange from NASA-Goddard

The stellar component of the Galaxy as seen by the AFGL I.R. Sky Survey 180Kb 810 stars

II/95/

B-I-M-N-T-f-v Grasdalen G.L., Gehrz R.D., Hackwell J.A., Castelaz M., Gullixson C., Astrophys. J. Suppl. Ser. 53, 413 (1983)

Id.: AFGL, IRC, HR, BD, other α, δ (1950) m_{v}, Sp, Variability Classification IK magnitudes AFGL magnitudes Observer

ANS UV Photometry Catalogue of Point Sources 476Kb 3573 sources

II/97/

B-I-M-N-T-f-v Wesselius P.R., van Duinen R.J., de Jonge A.R.W., Aalders J.W.G., Luinge W., Wildeman K.J., Astron. Astrophys. Suppl. Ser. 49, 427 (1982)

Id.: HD, DM α, δ Sp, m_{v}, UBV m and σ for 15N, 15, 18, 22, 25, 33 By exchange from NASA-Goddard Replaces II/60

A CAT-ALOGUE OF Hβ PHOTOM-ETRY OF SOUTHERN A5 TO G0 STARS BRIGHTER THAN 8.3 127Kb 2699 records

II/100/

B-I-M-N-T OLSEN E.H., PERRY C.L.: 1984, Astron. Astrophys. Suppl. Ser. 56, 229 (Tables II & III)

Id.: HD, DM β, σ, weight, N(obs)

Catalogue  References  Contents  Remarks

A CATALOGUE OF PHOTO-ELECTRIC MAGNITUDES AND COLOURS OF VISUAL DOUBLE AND MULTIPLE SYSTEMS 151Kb 2275 records

II/101/

B-I-M-N-T WALLENQUIST A.: 1981, Uppsala Univ. Astronom. Obs. Report 22 Id.: ADS, HD, BD, other α, δ (2000); Separation UBV, Sp Keypunched at the CDS

PRELIMINARY CATALOGUE OF PHOTO-GRAPHIC UBV PHOTOME-TRY FOR STARS IN OPEN CLUSTERS 1.7Mb 42575 stars

II/104/

B-I-L-M-N-T MERMILLIOD J.C.: 1984, Bull. Inf. CDS 27, 141 Id.: Geneva System V, B − V, U − B References

A Search for UV-excess objects 101Kb 1186 objects

II/106/

f-v Noguchi T., Maehara H., Kondo M., Ann. Tokyo Astron. Obs. 2nd Ser. 18, 55 (1980) Kondo M., Noguchi T., Maehara H. , Ann. Tokyo Astron. Obs. 2nd Ser, 20, 130 (1984)

Ident. : KUV name, field, other α, δ (1950) G-mag, U − G index Object code The two files of II/106/ were merged. by exchange from NASA-Goddard

THE 2-MICRON SKY SUR-VEY : THE NEAREST SAO STARS AND LOCATIONS ON POSS PRINTS 339Kb 6817 records

II/108/

B-I-M-N-T-f-v NAGY T.A. et al.: 1983, Astron. Data Center (ADC, NASA-Goddard) Bull. 1, 183 Id.: IRC, SAO Relative pos. IRC-SAO POSS , x, y modified Luyten Palomar number By exchange from NASA-Goddard

VILNIUS UP XY ZV T S PHO-TOMETRY 476Kb 2596 stars

II/109/

B-I-M-N-T NORTH P.: 1984, Bull. Inform. CDS 27, 133 Id.: Geneva System m_{v}, UP XY ZV T S N(meas.) References

Supersedes II/58/

A CATALOGUE OF UV IN-TERSTELLAR EX-TINCTION EXCESSES FOR 1415 STARS 191Kb 1415 records

II/114/

T-f-v SAVAGE B.D., MASSA D., MEADE M., WESSELIUS P.R.: 1985, Astrophys. J. Suppl. 59, 397 Id.: HD, BD, other α, δ (1950); ℓ, b Distance V, Sp E(B − V), E(33 − V), E(22 − V), E(18 − V), E(15 − V), E(Bump) Deviation, Remark By exchange from NASA-Goddard

Faint Blue Objects at high Galactic Latitude 234Kb 3987 sources

II/115A/

B-I-M-N-T-f-v Warnock A., Usher P.D., ADC Bull. 1, 195 (1993); updates: 1984, 2007 Id.: US, SA (57, 29, 28, 55, 71, 82, 94) α, δ (1950) m_{B}, color class Notes and Remarks By exchange from NASA-Goddard supersedes II/81/ SA 28, 29, 55, 57, 71, 82,

Catalogue  References  Contents  Remarks

PHOTOELECTRIC PHOTO-METRIC CAT-ALOGUE IN THE JOHNSON UBV RI SYSTEM 1.3Mb 6849 stars

II/116/

B-I-M-N-T-f-v LANZ T.: 1986, Astron. Astrophys. Suppl. Ser. 65, 195 Id.: Geneva system UBV RI N(meas.), Remarks References

CONCENTRIC APERTURE UBV RI PHOTOELECTRIC PHOTOMETRY OF GLOBU-LAR CLUSTERS 268Kb 117 clusters

II/117/

T-f-v PETERSON C.J.: 1986, Publ. Astron. Soc. Pacific 98, 192 Id.: GLC, IC, NGC, other Aperture UBV RI photometry References By exchange from NASA-Goddard

UBV RI PHOTOMETRIC STANDARDS AROUND THE CELESTIAL EQUATOR 85Kb 223 records

II/118/

B-I-N-T-f-v LANDOLT A.U.: 1983,, Astron. J. 88, 439 Id.: HD, SAO, Feige, other α, δ (1985) UBV RI N(obs), notes By exchange from NASA-Goddard

STELLAR DISTRIBUTION NEAR THE SOUTH GALACTIC POLE 35Kb 640 stars

II/119/

N-T-f-v BOK B.J., BASINSKI J.: 1964, Mem. Mount Stromlo Observatory 16 Id.: Number, Bok α, δ (1950) m_{pg}, (B − V)_{pg} by exchange from NASA-Goddard

A PHOTOMETRIC AND SPECTROPHOTOMET-RIC INVESTIGATION IN A REGION AT THE SOUTH GALACTIC POLE 149Kb 3312 records

II/120/

N-T-f-v ERIKSSON P-I.W.: 1978, Uppsala Astron. Obs. Report 11 Id.: Seq, HD/CD α, δ (1950) UBV for 477 stars BV_{pg} for 2835 stars by exchange from NASA-Goddard

STELLAR DISTRI-BUTION NEAR SELECTED AREAS 127, 141, and 189 2.6Mb 80486 stars

II/121/

N-T-f-v RATNATUNGA K.U., Ph.D. Dissertation, The Australian National University, Canberra (1983) α, δ (1950) J and F magnitudes by exchange from NASA-Goddard

UBV photoelectric photometry catalogue. I. The original data 13Mb 87263 stars

II/122B/

M-N-T-f-v Mermilliod J.C., Astron. Astrophys. Suppl. Ser. 71, 413 (1987)

Id.: Geneva System

UBV N(meas.) Reference See also II/193/

Catalogue  References  Contents  Remarks

UBV PHotometry of faint stars (V > 14.5) in the open cluster M67 68Kb 1117 records

II/123A/

B-I-M-N-T-f-v Frolov V.N., Ananyevskaya J.K. , Astron. Tsirk. 1432 (1986)

Id.: Number, Sanders UBV (V > 14.5) μ(relative) (from Sanders) Membership probability (from Sanders)

By exchange from Moscow Data Center see also V/52/

Catalogue of UBV Photometry and MK Spectral Types in Open Clusters (3rd Edition, August 1986) 2.0Mb 16087 stars

II/124A/

T-f-v MERMILLIOD J.C.: 1986, Astron. Astrophys. Suppl. Ser. 24, 159 (1st Ed.)

Id.: Geneva System, HD, DM V, B − V, U − B, Sp References Supersedes II/29/ and II/105/

IRAS Catalogue of Point Sources (V2.0, 1986) 45Mb 245889 sources

II/125/

B-I-L-M-N-T-f-v Joint IRAS Science W.G. , IPAC (1986)

Id.: IRAS α, δ (1950), uncertainty F (12, 25, 60, 100 μm) N(obs), qualities Association with known nearby sources By exchange from NASA-Goddard supersedes II/102/

IRAS Serendipitous Survey Catalog 8.8Mb 43886 sources

II/126/

T-f-v Kleinmann S.G., Cutri R.M., Young E.T., Low F.J., Gillett F.C., Joint IRAS Science W.G. (1986)

Id.: IRAS, field name, grid, ℓ, b α, δ F (12, 25, 60, 100μm) N(obs), qualities Association with known nearby sources By exchange from NASA-Goddard

Astrophysical data of visual double stars with early-type primaries 72Kb 544 components

II/127/

T-f-v Lindroos K.P., Astron. Astrophys. Suppl. Ser. 60, 183 (table VI) (1985)

Id. : HD Sp, E(b − y), M_{v} log L, log T_{eff}, age, Mass Distance, Separation Notes Keypunched at the CDS

PHOTOMETRIC DATA FOR THE BRIGHT STARS 353Kb 9094 records

II/128/

T-f-v HAUCK B., MERMILLIOD M.: 1986, Bull.Inf. CDS 31, 131 Id.: HR, HD m_{v} Photometric systems in which the star is measured Subset of II/167/

AUTOMATED 2-DIMENSIONAL CLAS-SIFICATION FROM MULTI-COLOUR PHOTOMETRY IN THE VILNIUS SYSTEM 75Kb 754 records

II/130/

B-I-M-T-f-v SMRIGLIO F., BOYLE R.P., STRAIZYS V., JANULIS R., NANDY K., McGILLIVRAY H.T., McLACHLAN A., COLUZZI R., SEGATO C.: 1986, Astron. Astrophys. Suppl. Ser. 66, 181 Id.: Number α, δ UP XY ZV T S Sp

Catalogue  References  Contents  Remarks

Tonantzintla Catalogue of the Pleiades Flare Stars 204Kb 519 stars

II/131/

N-f-v Haro G., Chavira E., Gonzalez G., Bol. del Instituto Tonantzintla, 3, N. 1 (1982)

Id.: Number, flare, Hertzsprung α, δ (1900) Flare-up observation dates U, B, V, m_{pg} at minimum, amplitudes Membership, Sp Notes, References

PHOTOGRAPHIC, PHOTO-VISUAL AND PHOTORED MAGNITUDES FOR 3124 STARS IN THE SE-LECTED AREA 40 131Kb 3197 records

II/132/

B-I-M-T KOLESNIK L.N.: 1987, Bull. Inform. CDS 33, 123 Id.: Number, BSD X, Y in BSD system m_{pg}, m_{v}, m_{R} By exchange from Moscow Data Center

PHOTOELECTRIC UBV OBSERVATIONS OF SUPERNOVAE 54Kb 1071 records

II/133/

B-I-M-N-T YOUNGER P.F., VAN DEN BERGH S.: 1985, Astron. Astrophys. Suppl. Ser. 61, 365 ; updated November 1987 Id.: SN, galaxy α, δ (1950) SN type, Hubble type Date of max. Blue extinction The catalogue is in print-out format

Catalog of Magnetic Field Measurements 153Kb 755 stars

II/134/

T-f-v Didelon P., Astron. Astrophys. Suppl. Ser. 53, 119 (1983)

Id.: HD / other α, δ (1950) m_{v}, Period, Sp H (magnetic field) Quality, Technique References

Update of the Catalogue of Diffuse Interstellar Bands by Snow et al. (1977) 96Kb 338 records

II/135/

T-f-v GUARINOS J., Bull. Inform. CDS 34 (1988)

Id.: HD, DM, other ℓ, b m, E(B − V), Sp W (eq. width), σ References Interstellar medium; Spectroscopy

A CATALOG OF BRIGHT uvbyβ STANDARD STARS 66Kb 366 stars

II/138/

B-N-T-f PERRY C.L., OLSEN E.H., CRAWFORD D.L.: 1987, Publ. Astron. Soc. Pacific 99, 1184 Id.: HR, IDS α, δ (2000) V, n_{V} , origins uvbyβ , origins References Statistics by exchange from NASA-Goddard

A Catalogue of the Brightest Stars in the field of M31 881Kb 11438 records

II/144/

B-I-M-N-T-f-v Berkhuijsen E. M., Humphreys R.M., Ghigo F.D., Zumach W., Astron. Astrophys. Suppl. Ser. 76, 65 (1988) Astron. Astrophys. 214, 68 Id.: Number (in δ-strip) α, δ (1950) UBV R (with σ)

See also II/145/

Catalogue  References  Contents  Remarks

A Catalogue of the Brightest stars in a foreground field of M31 688Kb 8778 stars

II/145/

B-I-M-N-T-f-v Berkhuijsen E. M., Humphreys R.M., Ghigo F.D., Zumach w., Astron. Astrophys. Suppl. Ser. 76, 65 (1988) Astron. Astrophys. 214, 68 (1989)

Id.: Number (in δ-strip) α, δ (1950) UBV R (with σ)

See also II/144/

A list of observations of the Ap star HR 1217 from 1980 to 1986 1.7Mb 52698 observations

II/146/

T-f-v Kurtz D.W. et al., Bull. Inform. CDS 36 (1989)

JD, m, filter Obs. reference

CHOPPED PHOTOMETRIC CHANNEL 41Mb 1508 observations

II/148/

N-T WESSELIUS P.R., BEIN-TEMA D.A., DE JONGE A.R.W., JURRIENS T.A. KESTER D.J.M., VAN VEERDEN J.E DE VRIES J., PERAULT M.: 1986 , Explanatory Supplement, Laboratory for Space Research Groningen 3 tar files including 3017 FITS files.

Each observation consists of two files: clean and raw images Each file consists of two maps (50 and 100 mum)

Catalogue of Diffuse Interstellar Band measurements 339Kb 2798 records

II/149/

T-f-v

Snow  T.P.,  York

D.G., Welty D.E., Astron. J. 82, 113 (1977) Guarinos J., Bull. Inform. CDS 35, 161 (1988)

Id.: HD, other

ℓ, b

m, Sp, E(B − V) DIB measurements Polarisation References Keypunched at the CDS

A Catalogue of Parameters for Eclipsing Binaries 130Kb 1048 binaries

II/150A/

N-T-f-v BRANCEWICZ H.K., DWORAK T.Z.: 1980, Acta Astron. 30, 501 Id.: Variable Name Sp, Classification Period, π, separation L, T_{eff}, Masses, Radius Roche-lobe filling up Keypunched at the CDS

OBSERVATIONS OF VARI-ABLE STARS 189Kb 8755 records

II/151/

B-I-M-N-T FEIJTH H., JURRIENS Th.: 1986, Id.: HV JD, m, Observer

INDIVIDUAL UBV OBSER-VATIONS OF STARS IN THE (LSS) CATALOG 73Kb 1732 stars

II/154/

N-T-f-v DRILLING J.S.: 1989, Louisiana State University Obs.

Id.: LSS Number Obs. date, Airmass UBV photometry by exchange from NASA-Goddard

Catalogue  References  Contents  Remarks

IRAS Faint Source Catalog, |b| > 10^{∘}, Version 2.0 52Mb 173044 sources

II/156A/

N-T-f-v Moshir M., Kopan G., Conrow T., McCallon H., Hacking P., Gregorich D., Rohrback G., Melnyk M., Rice W., Fullmer M., White J., Chester T.:, Explanatory Supplement to the IRAS Faint Source Survey, Version 2, California Inst. of Techn. JPL D-10015 (1992); Selected Astronomical Catalogues 1, Astronomical Data Center, NASA Id.: IAU α, δ (1950) Uncertainty ellipses F (12, 25, 60, 100μm) N(obs), Qualities Associations by exchange from NASA-Goddard

CATALOGUE OF PHOTO-ELECTRIC PHOTOMETRY IN THE VILNIUS SYSTEM, 1989. 766Kb 4849 stars

II/157A/

N-T-f-v STRAIZYS V., KAZLAUSKAS A., JODIN-SKIENE E., BARTKEVI-CIUS A.: , Astronomical Observatory of the Vilnius State University Id.: HD, other α, δ (1950) UP XY ZV S data Sp By exchange from Moscow Data Center Supersedes II/93/

THEORETICAL COLOURS AND ISOCHRONES FOR SOME HUBBLE SPACE TELESCOPE COLOUR SYS-TEMS 177Kb 279 + 650 records

II/160/

N-T EDVARSSON B., BELL R.A.: 1989 , Mon. Not. R. Astron. Soc. 238, 1121 Grid of synthetic surface brightness magnitudes for 21 HST WFC (Wide Field Camera) bandpasses by exchange from NASA-Goddard

EQUATORIAL INFRARED CATALOG 1.3Mb 7220 sources

II/161/

N-T-f-v SWEENEY L.H., RICHARDSON T.: 1989, Space Applications Corporation Id.: EIC, IRC, AFGL, AGK3, SAO, HD, BD α, δ (1950) + σ ; ℓ, b F (2.7μm), σ Distances to nearest stars Supersedes II/96/ by exchange from NASA-Goddard

A scanner CCD search for Cepheids and other variables in Crux and Centaurus. 121Mb 300308 stars

II/162/

f-v Caldwell J.A.R., Keane M.J., Schechter P.L. , Astron. J. 101, 1763 (1991)

Id.: Image+Number

α, δ (B1950)

Date, I_{c} at 7 epochs

Object class

CATALOGUE OF WALRAVEN PHOTOME-TRY 88 1.3Mb 5564 records

II/163/

B-I-M-N-T NITSCHELM C., MERMILLIOD J.C.: 1989, Astron. Astrophys. Suppl. Ser. 82, 331 Id.: LID, HR, HD, DM, other α, δ (1950) UBV W L indices References

Catalogue of DDO Photometry 88 1.5Mb 6139 stars

II/164/

B-I-M-N-T-f-v Mermilliod J.C., Nitschelm C., Astron. Astrophys. Suppl. Ser. 81, 401 (1989)

Id.: LID, HR, HD, DM, other α, δ (1950) m(35, 38, 41, 42, 45, 48, 51)

References

Catalogue  References  Contents  Remarks

CATALOGUE OF EGGEN’S uvby PHOTOMETRY 89 1.0Mb 4468 records

II/165/

B-I-M-N-T NITSCHELM C., MERMILLIOD J.C.: 1989 , Astron. Astrophys. Suppl. Ser. 85, 839 Id.: LID, HR, HD, DM, other α, δ (1950) V, b − y, m_{1}, c_{1} References

CATALOGUE OF WASHINGTON PHOTOME-TRY 89 373Kb 1000 records

II/166/

B-I-M-N-T MERMILLIOD J.C., NITSCHELM C.: 1989, Astron. Astrophys. Suppl. Ser. 84, 133 Id.: LID, HR, HD, DM, other α, δ (1950) V, (C, M, T_{1}, T_{2}, V, 51) References

THE GENERAL CATALOGUE OF PHOTOMETRIC DATA 6.2Mb 165896 stars

II/167/

B-N-T HAUCK B., NITSCHELM C., MERMILLIOD M., MERMILLIOD J.C.: 1990 , Astron. Astrophys. Suppl. Ser. 85, 989 Id.: LID (Lausanne Id) α, δ (1950) V Keys of photometric systems

Catalogue of homogeneous means in the UBV system 6.5Mb 92964 stars

II/168/

B-N-T-f-v Mermilliod J.C., (to be published, 191)

Id.: Geneva System V, B − V, U − B σ, N(obs) on CD-ROM: photom. and ubvmeas

Catalogue of stars measured in the Geneva Observatory Photometric System, 4^{th} Edition 5.7Mb 14633 stars

II/169/

B-I-N-T-f-v Rufener F., Geneva Observatory (1988)

Id.: Name, BS, HD, DM α, δ (1900) ; ℓ, b Membership to a cluster Normalized colours d, Δ, g, m_{2} indices V, σ_{V} , Sp, remarks

LONG-TERM PHO-TOMETRY OF VARIABLES AT ESO. I. 2.4Mb 534 stars

II/170/

B-N-T-f-v MANFROID J., et al.: 1991, Astron. Astrophys., Suppl. Ser. 87, 481 ESO Sci. Report 8 Id.: Code, HD, DM, other α, δ (2000), Sp JD, uvby Acquisition and Release codes

See also II/188/

Machine-Readable version of the Parenago Catalogue of stars in the area of the Orion Nebula 117Kb 2983 stars

II/171/

T-f-v Parenago P.P.: 1954 , Trudy Gos. Astron. Inst. im. Shternberga, 25 Malkov O.Yu.: 1992 , Bull. Inf. CDS 40, 13 Id.: Parenago, Variable name α, δ (1900) m_{pg}, m_{pv}, Sp by exchange from the Moscow Data Center

Catalogue  References  Contents  Remarks

A REFERENCE CATA-LOGUE OF GALACTIC NO-VAE 76Kb 283 novae

II/173/

B-N-T-f DÜRBECK H.W.: 1988 , Bull. Inf. CDS 34, 127 Id.: IAU, GCVS, other α, δ (1950); ℓ, b m(outburst), m(min) Light Curve Type Time of Decline Nova type Complete to the end of 1990

Catalogue of 166 ultraviolet extinction curves 307Kb 166 stars

II/175/

B-N-T-f-v Krelowski J., Papaj J., Wegner W., Bull. CDS 41, 23 (1992)

Id.: HD α, δ (2000) ; ℓ, b m_{V} , B − V, E(B − V) Sp, sources {λ^{−1}, E_{λ−2740}/E_{B−V} , σ}

NEAR-INFRARED PHOTOMETRY OF A SAMPLE OF IRAS POINT SOURCES 41Kb 516 sources

II/177/

B-N-T-f-v FOUQUÉ P., LE BERTRE T., EPCHTEIN N., GUGLIELMO F., KER-SCHBAUM F.: 1992 , Astron. Astrophys. Suppl. Ser. 93, 151 Id.: IRAS, optical association JHKLM photometry LRS class VH class

A CATALOGUE OF LINEAR POLARIZA-TION MEASUREMENTS for 5070 STARS 420Kb 5070 stars

II/178/

T-f-v AXON D.J., ELLIS R.S., Mon. Not. R. Astron. Soc. 177, 499 (1976)

Id.: HD or other ℓ, b Stokes parameters Q, U Distance, remarks by exchange from NASA-Goddard See also II/34/

Southern Spectrophotometric Standards 348Kb 62 stars

II/179/

T-f-v

Hamuy  M.,  Walker

A.R., Suntzeff N.B., Gigoux P., Heathcote S.R., Phillips M.M. , Publ. Astron. Soc. Pacific 104, 533 (1992) Publ. Astron. Soc. Pacific 106, 566 (1994) m(λ) in range λ 3300-7550Å by exchange from NASA-Goddard

uvbyβ PHOTOMETRY OF HIGH-VELOCITY AND METAL-POOR STARS 20Kb 220 stars

II/180/

B-N-T-f-v SCHUSTER W.J., NISSEN P.E.: 1989 , Astron. Astrophys. 222, 69-81 Id.: HD / DM / Giclas ... α, δ (1950) E(b − y), [F e/H] Age, σ(age) Notes

ATLAS AND CATALOGUE OF IN-FRARED SOURCES IN THE MAGELLANIC CLOUDS 408Kb 2140 sources

II/181/

N-T-f-v SCHWERING P.B.W., IS-RAEL F.P.: 1990 , Kluwer Academic Publishers, Dordrecht/Boston/London.

Id.: Number, IRAS α, δ (1950) F (12, 25, 60, 100 μm) + background Classification Cross-identifications with sources observed in other wavelengths by exchange from NASA-Goddard

Catalogue  References  Contents  Remarks

UBV PHOTOMETRY OF STARS WHOSE POSI-TIONS ARE ACCURATELY KNOWN 122Kb 3392 stars

II/182/

N-T-f-v OJA T.: 1984-1993 , Astron. Astrophys. Suppl. Ser. 57, 357 (1984); 59, 461 (1985); 61, 331 (1985); 65, 405 (1986); 68, 211 (1987); 71, 561 (1987); 89, 415 (1991); 100, 591 (1993); Id.: HR, BD, CoD UBV, N_{obs} Remarks

UBV RI Photometric Standard Stars in the magnitude range 11.5 < V < 16.0 around the Celestial Equator 101Kb 526 stars

II/183A/

T-f-v Landolt A.U. , Astron. J. 104, 340 (1992)

Id.: Feige, Giclas, PG, SA α, δ (J2000) UBV RI + σ N(obs), N(nights) Filter transmissions PM response

PHOTOGRAPHIC UBV PHOTOMETRY STAR CATALOGUE OF THE OLD OPEN CLUSTER NGC 188 166Kb 2171 stars

II/185/

M-N-T-f FROLOV V.N.: 1990 , Astron. Tsirk., 1544, 39 Id.: Number, Sandage x, y ; μ_{x}, μ_{y} UBV Membership probability Notes by exchange from the Moscow Data Center

30Dor IN THE LARGE MAG-ELLANIC CLOUD: THE STELLAR CONTENT AND INITAL MASS FUNCTION 282Kb 2395 stars

II/187/

N-T-f-v PARKER J.W.M.: 1992, Astron. J. 106, 560 Astron. J. 106, 1471 Publ. Astron. Soc. Pacific 104, 1107 Id.: Number α, δ (1950) ; X, Y UBV Sp, M_{v}, T_{eff}, M_{bol} Comments by exchange from NASA-Goddard

LONG-TERM PHO-TOMETRY OF VARIABLES AT ESO. II. THE SECOND DATA CATALOGUE (1986-1990) 2.9Mb 527 stars

II/188/

N-T-f-v STERKEN C., MANFROID J., et al.: 1993, Astron. Astrophys. Suppl. Ser. 102, 79 ESO Sci. Report 12 Id.: Code, HR, HD, DM, other JD, uvby Acquisition and Release codes

See also II/170/

SUPERNOVA STUDIES: A CATALOGUE OF MAGNI-TUDE OBSERVATIONS OF TYPE I SUPERNOVAE 313Kb 81 SNs

II/189/

T-f-v CADONAU R., LEI-BUNDGUT B.: 1990, Astron. Astrophys. Suppl. Ser. 82, 145 Id.: SN, host galaxy Host Galaxy parameters Date, mag., band, ref

Catalogue  References  Contents  Remarks

The IRAS Minor Planet Survey data base, 1992 7.7Mb 7311 observations

II/190/

N-T-f-v Tedesco E.F. (ed.): 1992, Phillips Laboratory Technical Report No. PL-TR-92-2049, Hanscom Air Force Base, MA Id.: Number, IAU H, albedo Obs. & Pred. α, δ ; λ, β Obs. & Pred. diameters and fluxes in IRAS bands by exchange from NASA-Goddard

UBV Photometry of Barium Stars 13Kb 109 stars

II/192/

T-f-v Watanabe E., Yutani M., Yamashita Y., Publ. Nat. Astron. Obs. Japan 3, 1 (1993)

Id.: HD or DM number α, δ (1900) V, B − V, U − B, their σ number of observation, group of Ba stars, Sp by exchange from NAO-Japan

UBV Photoelectric Catalogue: 1986-1992 Data 1.4Mb 25639 measurements

II/193/

N-T-f-v Mermilliod J.-C., Bull. Inf. CDS 45, 3 (1994)

Id.: Geneva system

UBV, ref Remarks

See also II/122/

A PHOTOMETRIC STUDY OF THE OPEN CLUSTER IC 4665 20Kb 200 stars

II/194A/

N-T-f-v McCARTHY M.F.S.J., O’SULLIVAN S.S.J.: 1969, Ric. Astron. 7, 483 Id.: Vasilevskis, Kopff, Alcaino V, B − V , σ Errors in the original publication corrected

INTERSTELLAR EX-TINCTION in the VICINITY of the NORTH AMERICA and PELICAN NEBULAE 239Kb 564 stars

II/196/

N-T-f-v

STRAIŽYS  V.,  KA-

ZLAUSKAS A., ČERNIS K., VANSEVIČIUS V: 1993, Baltic Astron. 2, 171 Id.: BD, HD, Metik α, δ (1950) Sp(lit., photom.) UP XY ZV S data

Interstellar Extinction in the direction of the open cluster IC 348 and the Per OB2 Association 99Kb 189 stars

II/197/

N-T-f-v Černis K.: 1993, Baltic Astron. 2, 214 Id.: Number, BD, HD α, δ (1950) Sp(lit., photom.) UP XY ZV S data

PHOTOELECTRIC PHOTOMETRY of CARBON STARS. 60Kb 178 stars

II/198/

T-f PAUPERS O., ŽAIME D., EGLĪTIS I.: 1993, Baltic Astron. 2, 268 Id.: CGCS, other UP XY ZV S data Obs. date

Catalogue  References  Contents  Remarks

Catalogue of Radio Stars 3.9Mb 3021 stars

II/199A/

N-T-f-v Wendker H.J., Astron. Astrophys. Suppl. Ser. 109, 177 (1995)

Id.: HD, DM, SAO

α, δ, μ m, Sp Radio Data References

Supersedes II/147/

LONG-TERM PHOTOM-ETRY OF VARIABLES AT ESO. III. THE THIRD DATA CATALOGUE (1990-1992) 770Kb 380 stars

II/200/

N-T-f-v MANFROID J., STERKEN C., CUNOW b., de GROOT M., JORISSEN A., KNEER R., KRENZIN R., KRUI-JSWIJK M., NAUMANN M., NIEHUES M., SCHOENE-ICH W., SEVENSTER M., VOS N., VOGT N. , Astron. Astrophys. Suppl. Ser. 109, 329 (1995) ESO Scientific Report 14 (1994)

Id.: Code, HR, HD, DM, other JD, uvby Acquisition and Release codes See also II/170/, II/188/, II/202/

Long-term photometry of variables at ESO. IV. (1992-1994) 925Kb 109 variables

II/202/

N-T-f-v Sterken C., Manfroid J., Beele D., de Koff S., Eggenkamp I.M.M.G., Göcking K., Jorissen A., Kaufer A., Kuss C., Schoenmakers A.P., Still J.M., van Loon J., Vink J., Vrielmann S., Wälde E. , Astron. Astrophys. Suppl. Ser. 113, 31 (1995) ESO Scientific Report 16 (1995)

Id.: Code, HR, HD, DM, other JD, uvby Acquisition and Release codes Corrected in October 1996 See also II/170/, II/188/, II/200/

Bibliographical Catalogue of Stellar Polarization 296Kb 1659 stars

II/206/

N-T-f-v Belous M.L., Bull. Inf. CDS 48, 5 (1996)

Id.: from literature

α, δ (2000) Sp, object type p, λ References

The Palomar-Green catalog of ultraviolet-excess stellar objects. 264Kb 1878 objects

II/207/

N-T-f-v Green R.F., Schmidt M., Liebert J., Astrophys. J. Suppl. Ser. 61, 305 (1986), Table 5 Id.: PG α, δ (1950); ℓ, b UBV, Class ugvr (AB_{79}) uvby Comments Table updated

BVRI CCD photometry of 361281 objects in the field of M 31 79Mb 485425 objects

II/208/

T-f-v Magnier E.A., Lewin W.H.G., van Paradijs J., Hasinger G., Jain A., Pietsch W., Trümper J., Astron. Astrophys. Suppl. Ser. 96, 379 Id.: Sequence number α, δ (J2000), σ BV RI + σ

Catalogue  References  Contents  Remarks

An imaging K-band survey 100Kb 634 sources

II/211/

T-f-v Glazebrook K., Peacock J.A., Collins C.A., Miller L. , Mon. Not. R. Astron. Soc. 266, 65 (1994) Mon. Not. R. Astron. Soc. 275, 169 (1995)

Id.: [GPC94], fieldID

α, δ (1950) KRBV I + σ z for 124 galaxies

OGLE Galactic Bulge periodic variables 235Kb 2861 variables

II/213/

T-f-v Udalski A., et al., Acta Astron. 44 317 (1994); 45, 1 (1995); 45, 433 (1995); 46, 51 (1996); 47, 1 (1997)

Id.: OGLE α, δ (J2000) P, Var. Type I_{max}, V − I, ΔI Remarks

Catalogue of uvbyβ Data 17Mb 63313 stars

II/215/

T-f-v Hauck B., Mermilliod M., Institut d’Astronomie de l’Université de Lausanne Id.: LID, HD, DM, other α, δ (1950) \underline{Mean Values}: V, b − y, m_{1}, c_{1}, β and σ Individual Measurements References

Supersedes II/158

Photoelectric observations of Cepheid variables in UBV(RI)c 4.8Mb 478 Cepehids

II/217/

f-v Berdnikov L.N., Sternberg Astronomical Institute, Moscow (1997)

Id.: Variable Name

α, δ (1900)

JD, (UBV R_{c}I_{c})

Reference

Polarisation of Be stars 57Kb 10 stars

II/220/

T-f-v McDavid D., Publ. Astron. Soc. Pacific 98, 572 (1986); Publ. Astron. Soc. Pacific 102, 773 (1990); Publ. Astron. Soc. Pacific 106, 949 (1994); Publ. Astron. Soc. Pacific 111, 494 (1999)

Obs. date, Passband Stokes parameters p, θ Precisions

From author

The Luminosity Function (LF) Catalogues 1.4Mb 18519 stars

II/221A/

f-v McCuskey S.W., Westerlund B., Seyfert C.K., Annear P.R., Nassau J.J., van Albada G.B., MacRae D.A., Farnsworth A.H., Astrophys. J. 109, 426 (1949); 109, 391 (1949); 110, 478 (1949); 112, 90 (1950); 118, 77 (1953); 130, 178 (1959); Astrophys. J. Suppl. Ser. 2, 75 (1955); 2, 123 (1955); 2, 271 (1956); 2, 298 (1956); 4, 1 (1959); 4, 23 (1959); 4, 73 (1959); Skiff B., (Unpublished, 1994-2003)

Id.: LF, GSC, DM, other α, δ (2000) recomputed m_{pg}, m_{i} Sp(1-dim) Remarks and Notes

Accurate Positions

Standard magnitudes in the E regions 123Kb 869 stars

II/222/

T-f-v Cousins A.W.J., Stoy R.H., R. Obs. Bull. 49, 3 (1962)

Id.: [CS62], HD, CPD

α, δ (1950)

U_{c}BV Photometry Notes, Remarks

Catalogue  References  Contents  Remarks

First Byurakan Survey (FBS), 2nd Program 56Kb 1103 objects

II/223/

T-f-v Abrahamian et al., Astrophysics 32, 14 (1990); 33, 418 (1990); 33, 493 (1990); 34, 7 (1991); 36, 62 (1993); 36, 306 (1993); 37, 27 (1994); 37, 117 (1994); 37, 224 (1994); 38, 108 (1995); 39, 315 (1996)

Id.: FBS α, δ (1950) m, spectral class Object type Existance in PG, KUV, CBS z Prepared by M. Vŕon (Obs. Haute-Provence, France)

Catalog of Apparent Diameters and Absolute Radii of Stars (CADARS) (3rd Edition) 1.6Mb 9733 stars

II/224/

T-f-v Pasinetti-Fracassini L.E., Pastori L., Pozzi A. , Dipartimento di Fisica, Universitá di Milano and Osservatorio Astronomico di Brera, Milano-Merate.

Id.: HD, BD, other m_{v}, Sp Apparent diameter Absolute radius Method of determination Source Remarks

Supersedes II/155

Catalog of Infrared Observations, Edition 5 34Mb 374653 obs.

II/225/

N-T-f-v Gezari D.Y., Pitts P.S., Schmitz M., Unpublished (1999)

Id.: IRAS, original α, δ (1950); beam size λ, F (λ), ref. IRAS flux code & quality by exchange from NASA-Goddard Supersedes II/216

An agglomeration of stellar polarization catalogs 1.9Mb 9286 stars

II/226/

T-f-v Heiles C., Astron. J. 119, 923 (2000)

Id.: HD, DM α, δ (J2000); (ℓ, b) Polarisation: %, θ, θ_{gal} m_{V} , Sp, Distance References

Supersedes II/34

The Asiago Supernova Catalogue 1999 239Kb 1447 SNe

II/227/

T-f-v Barbon R., Buondi V., Cappellaro E., Turatto M., Astron. Astrophys. Suppl. Ser. 139, 531 (1999)

Id.: Name, parent galaxy α, δ (J2000) Parent Galaxy parameters distance from the nucleus SN type m(max), date

Supersedes II/159

The DENIS Point Source Catalogue towards the Magellanic Clouds 335Mb 1635680 stars

II/228A/

T-f-v Cioni M.-R., Loup C., Habing H.J., et al., Astron. Astrophys. Suppl. Ser. 144, 235 (2000)

Id.: DCMC α, δ, σ (J2000) IJK photometry, BR(pg) σ(photometry) Flags in each band Qualities of Observations

UBV Phjotometry of Sakurai’s Object 10Kb 111 observations

II/229/

T-f-v Arkhipova V.P., et al. , Pis’ma Astron. Zh. 23, 709 (1997); 24, 297 (1998); 25, 711 (1999); Obs. JD UBV RI Photometer

Catalogue  References  Contents  Remarks

TASS Mark III photometric survey of the celestial equator 36Mb 367241 stars

II/230/

T-f-v Richmond M.W., Droege T.F., Gombert G., Gutzwiller M., Henden A.A.,i Albertson C., Beser N., Molhant N., Johnson H. , Publ. Astron. Soc. Pacific 112, 397 (2000)

Id.: Mark-III, AC2000, IRAS α, δ (J2000) V RI + σ Flags

Large-scale star formation in the Magellanic Clouds derived from analysis of stellar populations 20Mb 293694 stars

II/232/

f-v Braun J.M. , PhD Dissertation, Bonn Univ. (2001)

Id.: FieldSeq α, δ (J2000) UBV + σ (8 fields in LMC and SMC

Classification of Tycho variables stars 17Mb 1061 variables

II/233/

f-v Piquard S. , PhD Dissertation, Strasbourg, 2001 Id.: TYC, HIP, GCVS α, δ (J2000) Type of Variability Periods (Known and Found) Amplitude of variation

The ESO imaging survey: the Colour Catalog 10Mb 84085 objects

II/234/

f-v ESO , ESO imaging survey team Id.: EIS α, δ (J2000) UBV RIJK + σ in the Vega system

The Photometric Magnitude Difference Catalog 1.6Mb 19589 values

II/235/

f-v Worley C.E., Mason B.D., Wycoff G.L. , Astron. J. 122, 3482 (2001)

Id.: WDS, dboule star name Δm, method, reference Remarks

UBVR Survey of the Magellanic clouds 23Mb 264510 stars

II/236/

f-v Massey P., Astrophys. J. Suppl. Ser. 141, 81 (2002)

Id.: M2002, other α, δ (J2000) UBV R + σ Comparisons with other pohotometries Very blue and Very red stars

Catalogue of Stellar Photometry in Johnson’s 11-color system 519Kb 3945 records

II/237/

f-v

Id.: HD, DM, other α, δ (J1950), ℓ, b UBV RIJKLMN + σ

Catalogue  References  Contents  Remarks

Dahlmark variables 72Kb 432

II/238/

f-v Dahlmark L.; Skiff B.A., Williams G.V. , IAU Inf. Bull. Var. Stars 2157, 2878, 3855, 4329, 4448– 4451, 4458, 4642, 4734, 4898, 5181 (1982– 2001)

Id.: LD, GCVS, GSC, other α, δ (J2000) Variable Type P, T, A Remarks

A search for faint blue stars in high galactic latitudes I–III 397Kb 9010 stars

II/239/

f-v Berger J., Fringant A.M. , Astron. Astrophys. Suppl. Ser. 28, 123 (1977); 39, 39 (1980); 58, 565 (1984); Id.: PB, other α, δ (1950) m_{B}, Sp Remarks and Notes Supersedes II/111 and III/64

Differential V photometry of V350 Peg 495Kb 16191 measures

II/242/

f-v Vidal-Sainz J., Garcia-Melendo E., Lampens P., Van Cauteren P., Wils P. , Comm. in Asteroseismology (2003)

HJD, magnitudes

Observer

The ISOGAL Point Source Catalogue 76Mb 106150 sources

II/243/

f-v Schuller F., Ganesh S., Messineo M., Moneti A., Blommeart J.A.D.L., Alard C., Aracil B., Miville-Deschenes M.-A., Omont A., Schultheis M., Simon G., Soive A., Testi L., Astron. Astrophys. 403, 955 (2003)

Id.: IGPSC DENIS α, δ (J2000) ; ℓ, b 7 and 15 μm magnitudes IJK (DENIS) photometry Field details Quality and Association flags Spurious sources 384 raster images 16 sq.deg. in Milky Way

The 2MASS All-Sky Catalog of Point Sources 406Kb 470,992,970 sources

II/246/

f-v Cutri R.M., Skrutskie M.F., Van Dyk S., Beichman C.A., Carpenter J.M., Chester T., Cambrésy L., Evans T., Fowler J., Gizis J., Howard E., Huchra J., Jarrett T., Kopan E.L., Kirkpatrick J.D., Light R.M, Marsh K.A., McCallon H., Schneider S., Stiening R., Sykes M., Weinberg M., Wheaton W.A., Wheelock S., Zacarias N. , University of Massachusetts and Infrared Processing and Analysis Center (IPAC)/ California Institute of Technology Id.: 2MASS α, δ (J2000) JHK_{s} photometry + σ qualities, χ^{2} Observation Date Optical association See also VII/233 Supersedes II/241

Variable Stars in the Magellanic Clouds from the MACHO Project 8.5Kb 21474 variables

II/247/

f-v the MACHO Team , MA-CHO Project (2001)

Id.: Field.Tile.Seqn

α, δ (J2000) < V >, < R > P_{r}, P_{b} Variable class Light Curve Fits Variables in Magellanic Clouds

An investigation of globular cluster M 13 67Kb 923 stars

II/248/

f-v Kadla Z.I. , Izv. Glav. Astron. Obs. 181, 93 (1966)

Id.: Cl* NGC 6205 KAD UBV photometry Proper motion

Catalogues available at CDS  ( January 19, 2008)  55

Catalogue  References  Contents  Remarks

The Wisconsin H-Alpha Mapper 48Mb 37565 points

II/249/

f-v Haffner L.M., Reynolds R.J., Tufte S.L., Madsen G.J., Jaehnig K., Percival J.W., Astrophys. J. Suppl. Ser. 149, 405 (2003) ℓ, b I_{Hα} + σ FITS file with I(V) (intensity vs velocity) for each point

Combined General Catalog of Variable Stars (2004 Ed.) 17Mb 76604 variables

II/250/

f-v Samus N.N., Durlevich O.V., et al. , Institute of Astronomy of Russian Academy of Sciences and Sternberg State Astronomical Institute of the Moscow State University Id.: variable name α, δ (2000); (ℓ, b) m(max), m(min) Variability type Epoch, Period, D Sp Bibliography Galactic and Extragactic Variables Supernovae Merged List of Suspected Variables

Supersedes II/214

A  catalogue  of

the Chandra Deep Field South with multi-colour classification and photometric redshifts from COMBO-17 1241Mb 63501 objects

II/253/

f-v Wolf C., Meisenheimer K., Kleinheinrich M., Borch A., Dye S., Gray M., Wisotzki L., Bell E.F., Rix H.-W., Cimatti A., Hasinger G., Szokoly G., Astron. Astrophys. 421, 913 (2004)

Id.: Running number α, δ (J2000) Dimensions, Class UBV RI photometry z_{phot} 5 deep images in UBV RI

The SWIRE ELAIS N1 Image Atlas and Source Catalogs 73Mb 187467 sources

II/255/

f-v Surace J.A., Shupe D.L., Fang F., et al., Spitzer Science Center (2004)

Id.: SWIRE α, δ (J2000) Fluxes at 3.6, 4.5, 5.8, 8.0, 24 μm + σ Optical photometry Ug′r′i′Z Sources at 70 and 160μm

Sternberg Astronomical Institute Supernova Catalogue (2004) 563Kb 2991 SNe

II/256/

f-v Tsvetkov D.Yu., Pavlyuk N.N., Bartunov O.S. , Pis’ma Astron. Zh., 30, 803 (2004)

Id.: SN, NGC/IC, other α, δ (J2000) m_{pg} Parameters of Parent Galaxy Supernovae discovered until January 6, 2004 Supersedes II/218

A Finding List of Faint UV-Bright Stars 87Kb 734 stars

II/257/

T-f-v Lanning H.H., Meakes M., Publ. Astron. Soc. Pacific 85, 70 (1973) ; 106, 38 (1994) ; 107, 751 (1995); 110, 586 (1998) ; 112, 251 (2000); 113, 1393 (2001); 116, 1039 (2004)

Id.: Lanning α, δ (2000) B, U − B Notes

Supersedes II/231

The Hubble Ultra Deep Field Catalog 3.6Mb 10179 sources

II/258/

f-v Space Telescope Science Institute, 2004, Id.: Seq (F775W & F850LP filters) α, δ (J2000), X, Y Ellipticity, θ BV iz AB magnitudes +σ and S/N

Catalogue  References  Contents  Remarks

Extreme red stars in Cygnus 35Kb 267 stars

II/260/

f-v Ackermann G., Zeitschr. Astrophys. 69, 130 (1968)

Id.: ; α, δ (J2000) ; σ

Positions added by B. Skiff

The Great Observatories Origins Deep Survey: Initial Results from Optical and Near-Infrared Imaging 255Mb 61647

II/261/

f-v Giavalisco M., et al., Astrophys. J. 600, L93 (2004)

Id.: IAU

α, δ (J2000); (x, y)

SExtractor parameters in bviz bands: Ellipticity, area, Kron radius, class, magnitudes, flux

Beijing-Arizona-Taiwan-Connecticut (BATC) DR1 118Mb 511842 sources

II/262/

f-v Xu ZHOU, Zhaoji JIANG, National Astronomical Observatories of China (1995-2005)

Id.: USNO-A2.0 α, δ (J2000) Magnitudes in 15 intermediate-band filters between 300 and 1000nm

from Bejing Obs.

The All Sky Automated Survey. The catalog of variable stars. 363Mb 50122 stars

II/264/

f-v Pojmanski G., Acta Astron., 52, 397 Pojmanski G., Acta Astron., 53, 341 Pojmanski G., Maciejewski G., Acta Astron., 54, 153 Pojmanski G., Maciejewski G., Acta Astron., 55, 97 Pojmanski G., Pilecki B., Szczygiel D. , Acta Astron., 55, 275 Id.: ASAS α, δ (J2000) V I photometry , Period, Class Light Curves

The QUEST1 200k light curve catalog 358Mb 198213 light curves

II/266/

f-v Rengstorf A.W., Mufson S.L., Andrews P., Honeycutt R.K., Vivas A.K., et al. , Astrophys. J. 617, 184 (2004)

Id.: QUEST1, QUEST

α, δ (J2000) N_{obs} in bvr_{1}r_{3} Individual measurements

The MSX-UVISI Point Source Catalog 5.2Mb 47318

II/269/

f-v Newcomer R.E., Murthy J., Henry R.C., Price S.D., Paxton L., Air Force Research Lab. Technical Report AFRL-VS-TR-2004-1056 (2004)

Id.: J2000-based, Simbad UV magnitudes at 2230, 2310, 2480Å BV magnitudes of counterpart

TCS-CAIN: A deep multicolor NIR survey of the Galactic plane. 1194Mb 9471045 sources

II/270/

f-v Cabrera-Lavers A., Garzon F., Hammersley P.L., Vicente B., Gonzalez-Fernandez C. , Astron. Astrophys. 453, 371 (2006)

Id.: based on field, frame, number; 2MASS α, δ (J2000); ℓ, b JHK N(images)

Catalogue  References  Contents  Remarks

TASS Mark IV Photometric Survey of the Northern Sky. 632Mb 4353670 stars

II/271A/

f-v Droege T.F., Richmond M.W., Sallman M., Publ. Astron. Soc. Pacific 118, 1666 (2006); 2007 (in press)

Id.: Mark-IV α, δ (J2000) V I + σ

The Guide Star Photometric Catalog, Version 2.4 73Mb 554007 records

II/272/

f-v Bucciarelli et al. , Astron. Astrophys. 368, 335 (2001)

Id.: GSPC-II α, δ (J2000) BV R See also II/143/

New variable stars discovered in the MISAO project. 47Kb 500 variables

II/273/

f-v Yoshida S., Kadota K., Kato T. , Bull. Inf. Variable Stars 4746, 4770, 4771, 4780, 4793, 4812 (1999); 4842, 4854 (2000)

Id.: Mis, other variable names α, δ (J2000) m(min, max) Variability class

IRAS Point Source Reject Catalog, Version 2.1 (1989) 56Mb 372774 sources

II/274/

f-v Infrared Astronomical Satellite (IRAS) Catalogs and Atlases , vol. 1, Explanatory Supplement, 1988, ed. C. Beichman, et al., NASA RP-1190 (Washington, DC: GPO)

Id.: IRAS-R

α, δ (1950), uncertainty F (12, 25, 60, 100 μm) Qualities See II/125, II/156, II/126 and II/275

The SDSS Photometric Catalog, Release 5 108Kb 218218198 sources

II/276/

f-v Adelman-McCarthy J.K. et al. , Astron. J. (in prep.)

Id.: SDSS α, δ (J2000) ugriz Star/Galaxy Petrosian radius

Supersedes II/267

UBVRI photometry of faint field stars 2.6Mb 34071 stars

II/277/

f-v Skiff B., Lowell Observatory (2007)

Id.: GSC, Yale, BPS, ... α, δ (J2000) UBV R_{C}I_{C} References See also II/118

OGLE Search for planetary and low-luminosity object transits. 19Kb 177 transits

II/278/

f-v Udalski A., Paczynski B., Zebrun K., Szymanski M., Kubiak M., Soszinski I., Szewczyk O., Wyrzykowski L., Pietrzynski G. , Acta Astron., 52, 1 (2002) Acta Astron., 52, 115 (2002) Acta Astron., 52, 317 (2002) Acta Astron., 53, 133 (2003) Acta Astron., 54, 313 Id.: OGLE-TR α, δ (J2000) P er, T 0 Imag, V − I

III. Spectroscopic Data

Catalogue  References  Contents  Remarks

PRELIMINARY GENERAL CATALOGUE OF EARLY-TYPE EMISSION STARS 223Kb 3216 records

III/2/

B-I-M-N-T BERTIAU F.C., McCARTHY M.F.: 1969, Richerche Astronomiche 7, 523 Id.: Vatican, BD/DM m, Sp α, δ (1900), ℓ, b

BIBLIOGRAPHY OF STELLAR RADIAL VELOC-ITIES 5.7Mb 44133 records

III/4/

I-L-M-N-T

( microfiches ) [Data in Simbad]

ABT H.A., BIGGS E.S.: 1972, Kitt Peak Nat. Obs. Printed by Latham, New York Id.: star name, HR/HD α, δ (1900) Average V_{r}, N(obs) Reference Comments Closing date of the survey: June 1970

CATALOGUE OF FAINT OB STARS BETWEEN CARINA AND CENTAURUS 19Kb 285 records

III/5/

I-M-N-T LYNGA G.: 1968, Arkiv før Astronomi, 5, 181 (Medd. Lund Obs. Ser. 1, 238)

Id.: Number, CPD/HD m_{pg}, Sp α, δ (1950) ℓ, b

STUDIES OF THE MILKY WAY FROM CENTAURUS TO NORMA, III. OB STARS 32Kb 484 records

III/6/

I-M-N-T LYNGA G.: 1964, Acta Universitatis Lundensis Sectio II, 15 (Medd. Lund Obs. Ser. II, 141)

Id.: Number, CPD/HD m, Sp α, δ (1950) ℓ, b

A SURVEY OF FAINT OB STARS IN CARINA 33Kb 456 records

III/7/

I-M-N-T GRAHAM J.A., LYNGA G.: 1965, Memoirs Mt Stromlo Obs. 18 Id.: Number, HD/CPD m_{pg} α, δ (1975)

F8-G2 STARS IN THE NORTH GALACTIC POLE REGION 42Kb 1127 records

III/9/

I-M-N-T UPGREN A.R.: 1963, Astron. J. 68, 194 Id.: Number, BD, HD α, δ (1950) m_{pg} Remarks

G5 and LATER STARS in the NORTH GALACTIC POLE REGION 200Kb 4027 records

III/10A/

N-T-f-v UPGREN A.R.: 1962, Astron. J. 67, 37 See Upgren, Contr. Van Vleck Obs. 2, 19, 1984 for finder charts Id.: NGP, BD, HD m_{pg}, Sp Region by exchange from NASA-Goddard

Catalogue  References  Contents  Remarks

F2 AND EARLIER STARS IN S.A. 28, 54, 106, 107 16Kb 454 records

III/11/

I-M-N-T UPGREN A.R., STARON R.T.: 1969 , Astrophys. J. 157, 327

{Id.: BD \atop m_{pg}, Sp}

F5 AND LATER STARS IN S.A. 28, 54, 106, 107 72Kb 2068 records

III/12/

I-M-N-T UPGREN A.R., STARON R.T.: 1970 , Astrophys. J. Suppl. Ser. 19, 367

{Id.: BD \atop m_{pg}, Sp}

VYSSOTSKY’S CATALOGUES 1950.0 45Kb 895 records

III/13/

T-f-v VYSSOTSKY A.N. et al.: 1943-1958 , Astrophys. J. 97, 381 Astrophys. J. 104, 234 Astrophys. J. 116, 117 Astron. J. 61, 201 Astron. J. 63, 211 Id. : BD/AC m_{v}, Sp α, δ (1855) μ_{α}, μ_{δ}

STARS F2 AND EARLIER IN NORTH GALACTIC POLE REGION 23Kb 601 records

III/14/

I-M-N-T SLETTEBAK A., STOCK J.: 1959 , Hamburger Sternwarte Veröff. 5

Id.: BD, HD m, Sp α, δ (1950)

Catalogue of early-type stars whose spectra have shown emission lines 833Kb 5326 stars

III/17B/

N-T-f-v

[Identifiers in Simbad]

Wackerling L.R., Mem. Roy. Astron. Soc. 73, 153 (1970)

Id.: HD, DM, GC, LS, others m_{B}, m_{V} , Sp α, δ (1900), (1950) ell, b

Catalogue of stellar spectra classified in the Morgan-Keenan system 2.6Mb 20986 classifications

III/18B/

T-f-v

( microfiches ) [Data in Simbad]

Jaschek C., Condé H., de Sierra A.C., Publ. La Plata Obs. (1964)

Id.: Number, HD/DM

m_{pg}, Sp α, δ (1900) Bibliography A second file provides a coded MK spectral classification.

General Catalogue of Stellar Radial Velocity 974Kb 15106 stars

III/21/

I-M-N-T

( microfiches ) [Data in Simbad]

Wilson R.F., Carnegie Institution of Washington Publ. 601, Washington D.C. (1953)

Id.: Number, GC ,HD

α, δ(1950) ; ℓ, b

m_{pg}, Sp

V_{r} See III/190

Catalogue  References  Contents  Remarks

ROTATION OF EVOLVING A AND F STARS 37Kb 580 records

III/22/

B-I-M-N-T-f-v DANZINGER I.J., FABER S.M.: 1972 , Astron. Astrophys. 18, 428

Id.: HR ; m_{1}, c_{1}, b − y ; V sin i

MK classification for OB stars

26Kb 464 stars

III/23B/

M-N-f-v Lesh J.R., Astrophys. J. Suppl. Ser. 17, 371 (1968)

Id.: HR Sp Remarks from NASA-Goddard

ABUNDANCES OF SODIUM, MAGNESIUM AND CALCIUM IN K-TYPE GIANT STARS. TABLE I 26Kb 311 records

III/25/

I-M-N-T PEAT D.W., PEMBER-TON A.C.: 1968 , Mon. Not. Roy. Astron. Soc. 140, 21-41 Id.: HD Na(D), Mg(B), Ca (triplet)

(U − B), (B − V), (R − I)

ABUNDANCES OF SODIUM MAGNESIUM AND CALCIUM IN K-TYPE GIANT STARS. TABLE II 7.5Kb 85 records

III/26/

I-M-N-T PEAT D.W., PEMBER-TON A.C.: 1968 , Mon. Not. Roy. Astron. Soc. 140, 21-41 Id.: HD [Na/H], [Mg/H], [Ca/H], [Mg/Ca]

SCANNER ABUNDANCE STUDIES II. LATE G AND K DWARFS IN THE SOLAR NEIGHBORHOOD TABLE 5: RAW DATA 63Kb 238 stars

III/27/

I-M-N-T TAYLOR B.J.: 1970, Astrophys. J. Suppl. Ser. 22, 177-248 Id.: HD Magnitudes at λλ 4037, 4040, 4100, 4200, 4227, 4232, 4300, 4364, 4384, 4474, 4500, 4715, 4788, 4900, 5000, 5175, 5300, 5854, 5892, 6110, 6180, 6386, 6564, 6620, 7000, 7100, 7400, 8400, 8662, 8800

SCANNER ABUNDANCE STUDIES II. LATE G AND K DWARFS IN THE SOLAR NEIGHBOR-HOOD. TABLE 5: BLOCK-ING FRACTIONS 23Kb 283 records

III/28/

I-M-N-T TAYLOR B.J.: 1970, Astrophys. J. Suppl. Ser. 22, 177-248 Id.: HD Indices: CN(4200), Ca I(4227), Ca I(4232) CH (4300), Fe I(4383), MgH(4788), Mg I + MgH(5175), Na I(5892), Hα(6564), Ca II(8662) T temperature

A CATALOGUE OF Hγ MEASURES 74Kb 1171 stars

III/29/

I-M-N-T

[Data in Simbad]

PETRIE R.M., CRAMPTON D., LEIR A. YOUNGER F.: 1973, Publ. Dom. Astrophys. Obs. XIV, 151 Id.: HD / BD / ADS α, δ (1900) UBV, Sp W γ, N(meas) Remarks

Catalogue  References  Contents  Remarks

Michigan Catalogue of Two-Dimentional Spectral Types for the HD stars. Vol.I −90^{∘} < δ < −53^{∘} 3.6Mb 36382 stars

III/31B/

I-L-M-N-T-f-v

[Data in Simbad]

Houk N., Cowley A.P., Ann Arbor, Univ. of Michigan (1975)

Id.: HD, CPD m, Sp α, δ (1900) ; ℓ, b Precession Zone −90^{∘} < δ < −53^{∘} See also III/51/, III/80/, III/133/

ULTRAVIOLET BRIGHT STARS SPECTROPHOTO-METRIC CATALOGUE 906Kb 1356 stars

III/39A/

B-I-M-N-T-f-v JAMAR C., MACAU-HERCOT D., MONFILS A., THOMPSON G.I., HOUZI-AUX L., WILSON R.: 1976, ESA Special Report 27 F (λ), range 1360 to 2540Å Observations carried out by the S2/68 Ultraviolet Sky Survey Telescope aboard the ESRO satellite, TD1, from 19 March 1972 to 31 October 1972 See also II/86/

A uniform edition of the Stockholm Southern Milky Way Survey 1.0Mb 13259 stars

III/40A/

T-f-v Andersen J.; Nordstroem B. ; Loden et al. , Astron. Astrophys. Suppl. Ser. 29, 257 Id.: HD, HDE, DM Sp, m_{B} α, δ (1950) ; ℓ, b Remarks This catalogue includes III/33/, III/34/ and III/35/

Catalogue of Selected Spectral Types in the MK System 3.5Mb 30361 stars

III/42/

B-I-L-M-N-T-f-v

( microfiches ) [Data in Simbad]

Jaschek M., Bull. Inform. CDS 15, 121 (1978)

Id.: HD, other α, δ (1900) Selected MK spectral type

Catalogue of Luminous Stars in the Southern Milky Way 565Kb 5132 stars

III/43/

M-N-T-f-v Stephenson C.B., Sanduleak N., Publ. Warner & Swasey Obs. 1 (1971)

Id.: LSS, HD, DM, others m_{pg}, Sp α, δ (1950)

See also III/76

AN ATLAS OF STELLAR SPECTRA 4.1Mb 50 stars

III/44/

M-N-T-f-v JOHNSON H.L.: 1977, Rev. Mex. Astron. Astrofis. 2, 71; 4, 3 Id.: name F (λ) for 8192 points between 316.4nm and 1000nm Resolution 2 to 4 Å

Infrared Spectra for 32 stars 863Kb 46 spectra

III/45/

T-f-v Johnson H.L., Mendez M.E., Astron. J. 75, 785 (1970)

F (σ) for σ between 2500 and 8000 cm^{−1} Resolution: 8 cm^{−1}

Catalogue  References  Contents  Remarks

Catalogue of Ap and Am stars 37Kb 1231 stars

III/46/

I-M-N-T-f Bertaud Ch., Floquet M., Astron. Astrophys. Suppl. Ser. 16, 71 (1974) Bidelman W.P., MacConnell D.J. , Astron. J. 78, 687 (1973)

Id.: HD, other

p or m characteristic

Catalogue of Stellar Radial Velocities 806Kb 7823 stars

III/47B/

M-T-f-v

( microfiches ) [Data in Simbad]

Evans D.S., IAU Symp. 30, 57 (ed. A.H. Batten & J.F. Heard, Academic Press, 1967); See Bull. Inform. CDS 15, 121 Id.: HD, other α, δ (1900) m, Sp Mean V_{r} Number of plates, Origin See III/190 Only stars between 0 and 21h

CATALOGUE OF SPECTROPHOTOMET-RIC SCANS 436Kb 937 records

III/48B/

N-T-f Breger M., Astrophys. J. Suppl. Ser. 32, 7 (1976)

Id.: HR B − V, Sp λ, F (λ) Reference

Michigan Catalogue of Two-Dimentional Spectral Types for the HD stars. Vol. 2 3.8Mb 30400 stars

III/51B/

I-L-M-N-T-f-v

[Data in Simbad]

Houk N., Dept. of Astronomy, Univ. of Michigan Ann Arbor, Michigan Id.: HD, CPD, CoD α, δ, precession m, Sp Remarks −53^{∘} < δ < −40^{∘} By exchange from NASA/Goddard See also III/31/, III/80/, III/133/

CATALOGUE OF STARS CLASSIFIED FROM THE ULTRAVIOLET LINE FEATURES OF THE S2/68 EXPERIMENT 164Kb 1900 records

III/53/

B-I-M-N-T

( microfiches )

CUCCHIARO A., JASCHEK M., JASCHEK C.: 1979, Bull. Inform. CDS 17, 93 Id.: HD, HR α, δ (1950.0) m Sp(MK), Sp(UV) Remarks The microfiche contains also an Atlas illustrating 52 UV spectra

RADIAL VELOCITIES IN OPEN CLUSTERS 494Kb 6177 records

III/55/

I-M-N-T-f-v MERMILLIOD J.C.: 1979, Bull. Inform. CDS 16, 2 Id.: Geneva System JD, V_{r}with p.e. Number of lines Dispersion & Technique Reference Hyades, Pleiades, Praesepe, α Per and Coma Ber See also III/97/

A Catalogue of chemical elements identified in peculiar stellar spectra Version 1 (1976) 186Kb 1756 records

III/58/

B-I-M-N-T Levato H., Hernandez E.B., Bull. Inform. CDS 16, 93 (1979)

Id.: HD/BD m_{v}, Sp, Ion References

Catalogue  References  Contents  Remarks

A Catalogue and Bibliography of Mn-Hg stars 122Kb 127 stars

III/59/

I-M-N-T-f-v Schneider H., Astron. Astrophys. Suppl. Ser. 44, 137 (1981)

See description in Bull. Inform. CDS 20, 113 (1981)

Catalog of Be stars 204Kb 1159 stars

III/67A/

I-M-N-T-f-v

( microfiche )

Jaschek M., Egret D. , IAU Symp. 98, 261 (1982)

Id.: HD, DM, name α, δ (1950); ℓ, b UBV Sp(MK) V_{r}, v sin i Notes

A LIST OF EARLY-TYPE CHEMICALLY PECU-LIAR STARS 377Kb 3710 records

III/68/

I-M-N-T-f

( microfiches )

EGRET D., JASCHEK M.: 1981 , Comptes Rendus Symp. Liège 23, 495 Id.: HD, DM, CSI α, δ (1950) ; ℓ, b UBV References Includes Heabnormal, Ap and Am stars. This list is a part of CGS (Bull. Inf. CDS 20, 36)

White Dwarf Catalogues (I and II) 564Kb 6546 stars

III/70/

T-f-v Luyten W.J., White Dwarfs I (1970), White Dwarfs II (1979) Univ. of Minnesota, Minneapolis Id.: LWD, name α, δ (1950) ; μ m_{R}, m_{pg} Color spectral class by exchange from NASA-Goddard

Faint Blue Stars in the region near the South Galactic Pole 338Kb 8746 stars

III/74/

f-v Haro G., Luyten W., Bol. Tonantzintla y Tacubaya 22, 37-117 (1962)

Id.: PHL ; α, δ (1950) ; m_{pg}, UBV Keypunched at

CDS

MK Classification for HD stars in the 25^{∘} < δ < 30^{∘} zone 103Kb 1003 stars

III/75B/

I-M-T-f-v Jensen K.S., Private Communication (1982)

Id.: HD Sp, quality Detailed Remarks Remarks have been added

Catalogue  References  Contents  Remarks

Luminous Stars in the Northern Milky Way 451Kb 7392 stars

III/76A/

T-f-v Hardorp J., Rohlfs K., Stock J., Slettebak A., Hamburg-Bergedorf I (1959) Stok J., Nassau J.J., Stephenson C.B. , Hamburg-Bergedorf II (1960) Hardorp J., Theile I. Voigt H.H. , Hamburg-Bergedorf III (1964) Nassau J.J., Stephenson C.B. , Hamburg-Bergedorf IV (1963) Hardorp J., Theile I. Voigt H.H. , Hamburg-Bergedorf V (1965) Nassau J.J., Stephenson C.B., Mac-Donnell D.J. , Hamburg-Bergedorf VI (1965)

Id.: LS, BD α, δ (1950) Sp, remarks By exchange from NASA-Goddard See also III/43

0.2Åresolution far-UV Spectra Measured with Copernicus 6.8Mb 60 records

III/77/

T-f-v Snow T.P., Jenkins E.B., Astrophys. J. Suppl. Ser. 33, 269 (1977)

Id.: name

I(λ), N(obs) for λ = 1000 ⋯ 1450Å with a step of 0.2Å By exchange from NASA-Goddard

MK Classification Catalogue 2.3Mb 36450 stars

III/78/

B-I-L-M-T-f-v

[Data in Simbad]

Kennedy P.M., Mt Stromlo & Siding Spring Observatories, Australia (1983)

Id.: HD, other α, δ (2000) Sp, UBV Notes

SPECTRUM OF γ Tau 3.4Mb 45549 records

III/79/

B-I-M-N-T APPELQUIST L., ANDER-SEN J., FISHER W.A., FLETCHER J.M., KJAER-GAARD P.: 1983, Astron. Astrophys. Suppl. Ser. 52, 237 Ident.: λ and I(λ) in ranges: 5185.986 to 6005.007 Å 6000.009 to 7004.982 Å 6999.996 to 8004.987 Å 7999.981 to 8700.000 Å 5185.986 to 8700.000 Å

Catalogue of Two-dimensional Spectral Types for the HD stars. Vol.3 3.8Mb 30314 stars

III/80/

I-L-M-N-T-f-v

[Data in Simbad]

Houk N., Michigan Catalogue of Two Dimensional Spectral Types for the HD, 3, Ann Arbor, Dept. of Astronomy, Univ. of Michigan Id.: HD, DM, SAO Sp m_{pg} α, δ (1900) , ℓ, b precession Notes −40^{∘} < δ < −26^{∘} By exchange from NASA-Goddard See also III/31/, III/51/, III/133/

IUE Low-Dispersion Spectra Flux Catalogue. I. Normal Stars 3.7Mb 229 spectra

III/83/

T-f-v Heck A., Egret D., Jaschek M., Jaschek C., Astron. Astrophys. Suppl. Ser. 57, 213 (1984)

Id.: HD

λ, F (λ) (1152 ⋯ 3200Å) (2 Å intervals)

Wavelength range: 1152 − 3200Å

Catalogue  References  Contents  Remarks

Catalog of galactic O stars 105Kb 664 records

III/84B/

N-T-f-v Cruz-Gonzales C., Recillas-Cruz E., Costero R., Peimbert M., Torres-Peimbert S., Rev. Mex. Astron. Astrof., 1, 211 Id.: Number, Goy, other α, δ (1975); ℓ, b V, B − V Sp, quality V_{r}, quality Distances (r, z) HII region ident. Note flag, Remarks By exchange from NASA-Goddard

Far-UV objective-prism spectrophotometry: Skylab experiment S/019 6.9Mb 494 spectra

III/86/

I-M-N-T-f-v Heinize K.G., Wray J.D., Parsons S.B., Benedict G.F. , NASA Reference Publ. 1031 (1979)

Id.: HD, other name Exposure data Flux adjustment factor Flux data Magnitudes by exchange from NASA-Goddard

Stellar Spectrophotometric Atlas 2.0Mb 175 stars

III/88/

N-T-f-v Gunn J.E., Stryker L.L. , Astrophys. J. Suppl. Ser. 52, 121 (1983)

Id.: Number, HR, other V, A_{v}, UBV RI Airmass, α, δ by exchange from NASA-Goddard

A Catalog of Stellar Spectrophotometric Data 756Kb 1134 spectra

III/89/

T-f-v Adelman S.J., Pyper D.M., Shore S.N., White R.E., Warren W.H., Astron. Astrophys. Suppl. Ser. 81, 221 (1989)

Id.: Name, HD, HR B − V, Sp, comments Bandwidth, reference N(scans), HJD m(5000Å) (λ, m(λ)) by exchange from NASA-Goddard

A Library of Stellar Spectra 5.1Mb 161 spectra

III/92/

I-L-M-N-T-f-v Jacoby G.H., Hunter D.A. Christian C.A., Astrophys. J. Suppl. Ser. 56, 257 (1984)

Id.: HD, BD, SAO Sp UBV, E(B − V) F (λ) in range 3510–7427Å by exchange from NASA-Goddard

A CAT-ALOG OF RADIAL VELOC-ITIES IN THE LARGE MAG-ELLANIC CLOUDS .

III/93/

M-N-T-v FEITZINGER J.V., WEISS G.: 1979 , Astron. Astrophys. Suppl. Ser. 37, 575 Id.: Number, other α, δ (1950) ; X, Y V_{r}(observed and mean value) microfiche only Previously under number VI/13

BIBLIOGRAPHY OF INDI-VIDUAL RADIAL VELOCI-TIES FOR STARS IN OPEN CLUSTERS. II. NGC AND IC CLUSTERS 524Kb 6626 stars

III/97A/

I-M-N-T-f MERMILLIOD J.C.: 1984, Bull. Inform. CDS 26, 9 Id.: Geneva System JD, V_{r}+ σ Number of lines Dispersion & Technique Reference See also III/55/

Catalogue  References  Contents  Remarks

RADIAL VELOCITIES OF SOUTHERN STARS OB-TAINED WITH THE PHO-TOELECTRIC SCANNER CORAVEL- III 109Kb 790 stars

III/105/

I-M-N-T ANDERSEN J., NORDSTROM B., ARDE-BERG A., BENTZ W., IM-BERT M., LINDGREN H., MARTIN N., MAURICE E., MAYOR M., PREVOT L.: 1985, Astron. Astrophys. Suppl. Ser. 59, 15

Id.: HD, HR m_{v}, Sp V_{r}, Remarks

MK Classification of Visual Multiples 119Kb 793 systems

III/106/

I-M-N-T-f-v Abt H.A., Astrophys. J. Suppl. Ser. 45, 437 (1981)

Id.: ADS, HD, HR

Selected published types Keypunched at the CDS

CATALOGUE, SPECTRUM AND MAGNITUDE DATA BANK OF B(e), B(p) and B(pe) STARS 3.2Mb 27575 records

III/110/

B-I-M-N-T PAGE A.A.: 1981, 1984, Mount Tamborine Obs. Publ. 2; version 2: 1984 Id.: name, HD, BD, CD, CPD, AGK, SAO, HR, GC, Yale, CPC, CZC, LS, MWC, Ton, Hiltner, Wray, Henize, Jenkins Sp, v sin i m_{v}, m_{pg}, B − V Variability References

Supersedes III/71/

Catalogue of Spectral and Luminosity classes of 10396 stars in Kapteyn areas 2-43 910Kb 10492 stars

III/112/

T-f-v Bartaya R.A., Bull. Abastumani Astrophys. Observ. 51 (1979) Bull. Inform. CDS 24, 93 (1983)

Id.: SA, BD, HD, SAO

α, δ (1950) m_{pg}, Sp(BSD) μ_{α}, μ_{δ} Comments By exchange from Moscow Data Center new version of III/72/

A DEEP OBJECTIVE-PRISM SURVEY FOR LARGE MAGELLANIC CLOUD MEMBERS 73Kb 1273 stars

III/113/

I-M-N-T SANDULEAK N.: 1969, Cerro Tololo Inter-American Obs. Contr.

Id.: Number, HD, CPD α, δ Sp, m_{pg}, variability References By exchange from NASA/Goddard supersedes III/94/

Spectra of late-type standard stars in the region 2.0-2.5 μm 1.1Mb 26 spectra

III/114/

B-I-M-N-T-f-v Kleinmann S.G., Hall D.N.B. , Astrophys. J. Suppl. Ser. 62, 501 (1986)

Id.: name, HR

Date λ, F (λ) by exchange from NASA-Goddard

IUE Atlas of O-type stellar spectra 6.5Mb 1200 spectra

III/115/

I-M-N-T-f-v Walborn N.R., Nichols-Bohlin J., Panek R.J., NASA Reference Publ. 1155 (1985)

Id.: HD/HDE, Flamsteed, other Sp SWP, investigator fluxes, splices By exchange from NASA-Goddard See also III/188

Catalogue  References  Contents  Remarks

Spectrophotometric Standards (1987) 1.6Mb 25 spectra

III/116/

T-f-v Massey P., Strobel K., Barnes J.V., Anderson E., Astrophys. J. 328, 315 (1988)

Id.: HR, Name N(points) λ (Å) of first observation step (Å) magnitudes By exchange from NASA-Goddard

Close Visual Binaries. I. MK Classifications 79Kb 170 binaries

III/117/

N-T-f-v Corbally C.J. , Astrophys. J. Suppl. Ser. 55, 657( 1984)

Id.: HR, HD, IDS

Sp, UBV, E(B − V), M_{v} Notes and remarks

G.P. KUIPER’S SPECTRAL CLAS-SIFICATIONS OF PROPER-MOTION STARS 147Kb 3172 records

III/119/

I-M-T BIDELMAN W.P.: 1985, Astrophys. J. Suppl. Ser. 59, 197 Id.: LTT, name m, Sp Notes Keypunched at the CDS

CATALOGUE OF STARS WITH CaII H AND K EMISSIONS 107Kb 1846 stars

III/120/

I-M-N-T-f-v GLEBOCKI R., MUSIELAK G.: 1980 , Acta Astron. 30, 453 Id.: HD/BD Sp, V, B − V I(H, K), W Ca Reference, Remarks Keypunched at the CDS

CATALOGUE OF STELLAR ABUNDANCES 1.0Mb 16716 records

III/121/

I-M-N-T-f MARECKI A., STROBEL J., STROBEL A.: 1986, Torun Observatory Id.: HD Ion, standard abundance Reference comments

Dwarf K and M stars of small proper motion found in a large spectroscopic survey 272Kb 2201 stars

III/123B/

T-f-v Stephenson C.B., Astron. J. 91, 144 (1986)

Id.: Number, name

α, δ (1900)

m, Sp

Additional remarks Keypunched at the CDS

Spectrophotometry of bright F, G, K and M-type stars. 60 Southern and Equatorial stars 246Kb 60 stars

III/124/

I-M-N-T-f-v Kiehling R., Astron. Astrophys. Suppl. Ser. 69, 465 (1987)

Id. : HR or HD

Spectral energy distribution in range 320–880 nm

Catalogue  References  Contents  Remarks

Catalogue of metal-deficient F-M stars classified spectroscopically. Supplement 1. (MD-SPS1) 144Kb 900 stars

III/125/

I-M-T-f-v Bartkevicius A., Bull Vilnius Obs. 68, 3 (1984) Bull Vilnius Obs. 74, 55 (1986)

Id.: Number, names α, δ (1950), ℓ, b UBV, color excess [F e/H], Sp By exchange from Moscow Data Center see also IV/20/

Stellar Spectrophotometric Catalogue 9.9Mb 1588 stars

III/126/

I-M-N-T-f-v Burnashev V.I., Abastumanskaya Astrofiz. Obs. Bull. 59, 83 (1985)

Id.: BS/HD, name m, Sp, N(obs) λ, log F (λ) (3200 to 7350 Å, step 25 Å) References By exchange from Moscow Data Center

RADIAL VELOC-ITIES OF BRIGHT SOUTH-ERN STARS 182Kb 743 stars

III/127/

I-M-N-T-f ANDERSEN J., NORD-STRÖM B.: 1983 , Astron. Astrophys. Suppl. Ser. 52, 471; 52, 479 NORDSTRÖM B., ANDERSEN J.: 1985 , Astron. Astrophys. Suppl. Ser. 61, 53 Id.: HD, HR, FK4 m_{v}, Sp V_{r}, H and K velocities Mean errors Rotation class Remarks By exchange from NASA-Goddard Supersedes III/102, III/103 and III/104

Atlas of the wavelength dependence of UV extinction in the Galaxy 75Kb 2258 records

III/130/

T-f-v Aiello S., Barsella B., Chlevicki G., Mayo Greenberg J., Patriarchi P., Perinotto M.: 1988, Astron. Astrophys. Suppl. Ser. 73, 195

{Id.: HD \atop λ, E(λ)/E(B − V)}

A CRITICAL CATALOG OF SPECTROSCOPIC STEL-LAR ABUNDANCE ANALY-SIS 1.4Mb 32109 records

III/131/

T-f-v KÖPPEN J.: 1987, Id.: HD, BD, cluster, other Sp, m_{v}, [F e/H] References, methods Remarks, comments Index Software

CATALOGUE OF TWO-DIMENSIONAL SPEC-TRAL TYPES FOR THE HD STARS. Vol.4 4.1Mb 33301 stars

III/133/

B-I-M-N-T-f-v HOUK N., SMITH-MOORE M.: 1988, Michigan Catalogue of Two Dimensional Spectral Types for the HD, 4, Ann Arbor, Dept. of Astronomy, Univ. of Michigan Id.: HD, DM, SAO Sp, m_{pg} α, δ (1900); ℓ, b Notes Zone −26^{∘} < δ < −12^{∘} By exchange from NASA-Goddard See also III/31/, III/51/, III/80/

Catalogue  References  Contents  Remarks

RADIAL VELOC-ITIES OF BRIGHT SOUTH-ERN STARS. V. 13Kb 146 records

III/134/

B-I-M-N-T ANDERSEN J., NORD-STROM B.: 1985, Astron. Astrophys. Suppl. Ser. 62, 355 Id.: HD Julian date Weight flag, num. lines V_{r}, σ(V_{r}) Rotation class Remarks By exchange from NASA-Goddard

The Henry Draper Catalogue (edition 1989) 12Mb 272150 stars

III/135A/

N-T-f-v Cannon A.J., Pickering E.C.: 1918-1925, Ann. Harvard College Obs. 91-100 Id.: HD, DM, AG α, δ (1900) m_{v}, m_{pg} Sp Estimated phot. intens. Remarks By exchange from NASA-Goddard Supersedes III/99/

An Atlas of optical spectrophotometry of Wolf-Rayet Carbon and Oxygen stars 3.0Mb 88 spectra

III/136/

B-I-M-N-T-f Torres A.V., Massey. P., Astrophys. J. Suppl. Ser. 65, 459 (1987)

Id.: HD, other name(s) Sp Absorption flag Instrument code Number of data points Comment Spectra By exchange from NASA-Goddard

A Catalogue of Spectroscopically Identified Hot Subdwarf Stars 175Kb 1225 stars

III/137/

f-v Kilkenny D., Heber U., Drilling J.S. , S. Afr. Astron. Obs. Circ. 12, 1 (1988)

Id.: HD, LS, G, GR, other α, δ (1950) UBV photometry, Sp T_{eff}, log g References

A Compendium of Equivalent Width Measures 1.7Mb 287 stars

III/139A/

N-T-f-v Luck E.R., Case Western Reserve University (18 publications in Astrophys. J. and Astrophys. J. Suppl. between 1977 and 1991)

Id.: Name, HD or HR Species, λ, χ, g_{f} , W Fortran program by exchange from NASA-Goddard

Cool Carbon Stars found with the Baldone Schmidt Telescope (BC) 33Kb 318 records

III/140/

T-f-v ALKSNE Z., ALKSNIS A., OZOLINA V., PLATAIS I.: 1987, Radioastrophys. Observatory Latvian Academy of Sciences, Riga Id.: Number α, δ ; ℓ, b Spfor 318 stars Photom. data (157 stars)

By exchange from Moscow Data Center Update of III/73/

RATES OF MASS LOSS 9.0Kb 274 records

III/141/

B-I-M-N-T DE JAGER C., NIEUWEN-HUIJZEN H., VAN DER HUCHT K.A.: 1988, Astron. Astrophys. Suppl. Ser. 72, 259 Sequence number T_{eff}, L, Ṁ weight, O − C Contains only columns 1, 5, 7, 9, 10 and 11 of the printed version

Catalogue  References  Contents  Remarks

Spectrophotometry of Wolf-Rayet Stars 5.4Mb 174 spectra

III/143/

B-I-N-T-f-v Torres-Dodgen A.V., Massey P., Astron. J. 96, 1076 (1988)

Id.: Name of star F (λ) Dispersion Supersedes most of III/136/ By exchange from NASA-Goddard

Radial Velocities for 889 high proper motion stars measured with the Mount Wilson 100 inch reflector 101Kb 878 stars

III/144/

B-I-M-N-T-f-v Fouts G., Sandage A., Astron. J. 91, 1189 (1986)

Id.: Name of star α, δ (1950) ; ℓ, b UBV μ_{α}, μ_{d}elta ; μ_{ℓ}, μ_{b} V_{r} N(obs)

Keypunched at the CDS

U, V, W Velocity Components for the Old Disk using Radial Velocities of 1295 Stars in the Three Cardinal Directions 63Kb 1295 stars

III/145/

f-v Sandage A., Fouts G. , Astron. J. 93, 923 (1987)

Id.: SAO, BD α, δ (B1950) BV magnitudes V_{r} + σ

STUDIES OF THE LMC STELLAR CONTENT 144Kb 1827 records

III/147/

T-f-v ROUSSEAU J., MARTIN N., PREVOT L., REBEIROT E., ROBIN A.: 1978, Astron. Astrophys. Suppl. Ser. 31, 243 Astron. Astrophys. Suppl. Ser. 31, 243 Id.: HD, R, NS, G, FDS, OM, C, D, BI, FDA Chart of HW atlas α, δ (1975) Sp, references m_{pg}, UBV

Supersedes III/111/

Interstellar Lines Catalogue 1.6Mb 2130 objects

III/148A/

T-f-v Garcia B., Astron. Astrophys. Suppl. Ser. 89, 469 (1991)

Id.: HD, DM α, δ (1950) magnitude, Sp, reference interstellar line measured and type of measure interstellar diffuse band equivalent width observations, fundamental data cross-identifications bibliography Keypunched at the CDS

A LIST OF MK STANDARD STARS 130Kb 963 stars

III/149/

N-T-f-v GARCIA B.: 1989, Bull. Inform. CDS 36, 27 Id.: HD/BD, HR, name α, δ (1950) Sp Special spectrum features m, B − V, references Keypunched at the CDS

The Perkins catalog of revised MK types for the cooler stars 216Kb 1137 stars

III/150A/

B-I-N-T-f-v Keenan P.C., McNeil R.C. , Astrophys. J. Suppl. Ser. 71, 245 (1989)

Id.: name, variable designation, Gliese, Giclas, HR, HD α, δ (2000) Standard star flag V or m_{pg}, uncertainty flag, variability flag Sp, remarks Supersedes III/66

Catalogue  References  Contents  Remarks

Catalogue of K Stars Towards the South Galactic Pole 159Kb 2016 stars

III/151/

f-v Kuijken K., Gilmore G. , Mon. Not. R. Astron. Soc. 239, 305 (1989)

Id.: Running number

α, δ (B1950)

UBV (photographic)

V_{LSR}

An Atlas of Stellar Spectra between 2.00 and 2.45 μm 165Kb 83 spectra

III/153/

B-I-N-T-f-v Arnaud K.A., Gilmore G., Cameron A.C., Mon. Not. R. Astr. Soc. 237, 495 (1989)

Id.: Header

λ, F (λ) in range 2.0 − −2.45μm Fractional error By exchange from NASA-Goddard

Radial Velocities in three fields along the Southern Galactic Equator 51Kb 764 stars

III/154/

B-M-N-T-f-v Denoyelle J. , Astron. Astrophys. Suppl. Ser. 70, 373 (1987)

Id.: CPD α, δ (1950) V_{r}, N(plates), σ

THE SECOND HENRY DRAPER EXTENSION : SPECTRAL TYPES, MAGNITUDES AND PRECISE POSITIONS FOR 10639 stars 670Kb 11282 records

III/155/

T-f CANNON A.J.: 1937 , Annals Astron. Obs. Harvard Coll. 105, 1 RÖSER S., BASTIAN U., WIESE K.: 1991, Astron. Astrophys. Suppl. Ser. 88, 277 Id.: HDE, PPM m_{pg}, Sp α, δ (2000.), epochs, μ

An Ultraviolet Atlas of Quasar and Blazar Spectra 1.4Mb 69 objects

III/157/

N-T-f-v Kinney A.L., Bohlin R.C., Blades J.C., York D.G., Astrophys. J. Suppl. Ser. 75, 645 (1991)

Id.: IAU system

λ, quality, F (λ), σ

CaII H and K measurements made at Mount Wilson Observatory, 1966-1983 494Kb 3111 measurements

III/159A/

N-T-f-v Duncan D.K., et al., Astrophys. J. Suppl. Ser. 76, 383 (1991)

Id.: HD, DM, other

S-index (min, max, mean, σ) Weights, dates by exchange from NASA-Goddard

General Catalogue of Ap and Am stars 989Kb 6684 stars

III/162A/

N-T-f-v Renson P., Gerbaldi M., Catalano F.A., Astron. Astrophys. Suppl. 89, 429 (1991)

Id.: Bertaud, HD, DM, GC, Other α, δ (1900, 1950) UBV, uvbyβ Sp, CP nature Notes about binarity, variability, magnetic field,...

About 12 modifications in the Notes

Catalogue  References  Contents  Remarks

FOURTH CATALOGUE OF Am STARS WITH KNOWN SPECTRAL TYPES 301Kb 2152 stars

III/163/

I-N-T HAUCK B.: 1992 , Bull. Inf. CDS 40, 19 Id.: Name, Cluster, HD, ... α, δ (1950) Sp(K), Sp(H), Sp(m), ref. V, V. sin i, ref.

Supersedes III/107/

A Master Catalogue of Equivalent Widths of the Interstellar 217nm band 169Kb 2861 measurements

III/164/

N-T-f-v Friedemann C. , Bull. Inf. CDS 40, 31 (1992)

Id.: HD / BD E(B − V), λ_{c} Parameters ABCW Number of obs., remarks

Supersedes III/95/

A New Library of Optical Spectra 432Kb 72 spectra

III/166/

T-f-v Silva D.R., Cornell M.E., Astrophys. J. Suppl. Ser. 81, 865 (1992)

Id.: HD, other Sp, T_{eff}, E(B − V) U − B and B − V synthetic colours Intensities at 5Å intervals in range 3510–8930Å by exchange from NASA-Goddard

AN ATLAS OF ULTRAVIO-LET SPECTRA OF STAR-FORMING GALAXIES 12Mb 143 galaxies & Spectra

III/167/

N-T-f KINNEY A. L., BOHLIN R. C., CALZETTI D., PANA-GIA N., WYSE R. F.: 1993 , Astrophys. J. Suppl. Ser. 86, 5 Id.: NGC, IC, Mrk, ... α, δ (1950); ℓ, b Morph.Type, Activity cl. v_{H}, B_{T}, M_{B} by exchange from NASA-Goddard

A General Catalog of S Stars, 2nd Ed. 252Kb 1347 stars

III/168/

N-T-f-v STEPHENSON, C.B.: 1984, Publ. Warner & Swasey Obs. 3, No.3, 1 Id.: Number (CSS), other α, δ (1900) ; ℓ, b BV I photographic mag. Sp Notes and references Rejected S stars by exchange from NASA-Goddard Supersedes III/60/

A Photometric Atlas of the Spectrum of Procyon (3140– 7470Å) 16Mb 866,000 points

III/169A/

f-v Griffin R.E., Griffin R., R. & R. Griffin, publishers, Cambridge, U.K.(1979)

I(λ), with Δλ =5mÅ for α CMi (HR 2943) by exchange from NASA-Goddard

IRAS Point Source Identifications 976Kb 14192 sources

III/170/

N-T-f-v MacConnell D. J., (unpublished) (1993)

Id.: IRAS Distance from IRAS source

Classification Plate identification Remarks by exchange from NASA-Goddard

RADIAL VELOCITIES FROM OBJECTIVE PRISM PLATES 102Kb 1302 stars

III/172/

N-T-f-v STOCK J.: 1992 , Rev. Mex. Astron. Astrofis. 24, 45 Id.: Number α, δ (1950), σ m_{pg}, Sp V_{r}, σ, N(plates)

Two areas: one near SGP, one in plane

Catalogue  References  Contents  Remarks

An Atlas of high resolution line profiles of Symbiotic stars 10Mb 436 spectra

III/174/

N-T-f-v van Winckel H., Duerbeck H. W., Schwarz H. E., Astron. Astrophys. Suppl. Ser. 102, 401 (1993) Ivison R.J., Bode M.F., Meaburn J., Astron. Astrophys. Suppl. Ser. 103, 201 (1994)

Id.: Variable, other

α, δ (1950) High-resolution spectra around 6563Å(Hα), 6673Å(HeI), 4686Å(HeII), 6725Å(SII), 5007A(OIII) Low-resolution spectra 82 symbiotic stars

Optical spectroscopy of 1Jy, S4 and S5 radio source identifications 54Kb 126 sources

III/175/

N-T-f-v Stickel M., Kuehr H., Fried J.W. , Astron. Astrophys. Suppl. Ser. 80, 103; 97, 483; 100, 395; 101, 521; 103, 349 Id.: Radio name α, δ (1950) S(5GHz) Optical type, z, Flux Obs. date, Tel.

RR Lyr DATA II. THE METALLICI-TIES AND KINEMATICS OF LOCAL RR LYRAE 31Kb 302 stars

III/176/

T-f LAYDEN A.C.: 1994, Astron. J. (in press)

Id.: GCVS

V, [Fe/H], with σ

V_{r}, with σ

Distance, galactic position by exchange from NASA-Goddard

SURVEY OBSERVATIONS OF EMISSION-LINE STARS IN THE ORION REGION 128Kb 1214 stars

III/177/

N-T-f-v WIRAMIHARDJA S.D., KOGURE T., YOSHIDA S., OGURA K., NAKANO M. : 1989 , Publ. Astron. Soc. Japan 41, 155 (Part I) KOGURE T., YOSHIDA S., WIRAMIHARDJA S.D., NAKANO M., IWATA T., OGURA K.: 1989, , Publ. Astron. Soc. Japan 41, 1195 (Part II) WIRAMIHARDJA S.D., KOGURE T., YOSHIDA S., NAKANO M., OGURA K., IWATA T.: 1991, Publ. Astron. Soc. Japan 43, 27 (Part III) WIRAMIHARDJA S.D., KOGURE T., OGURA K., YOSHIDA S., NAKANO M., : 1993 , Publ. Astron. Soc. Japan 45, 643 (Part IV)

Id.: area + serial number; previous works α, δ (1950) Hα emission intensities, V magnitude ID with the previous surveys by exchange from NAO-Japan

An atlas of Southern MK standards from 5800 to 10200 Å 675Kb 137 spectra

III/179/

f-v Danks A.C., Dennefeld M., Publ. Astron. Soc. Pacific 106, 382 (1994)

Id.: HD α, δ (J2000) Sp, (BV) Spectrum

Catalogue  References  Contents  Remarks

New H-alpha Emission Stars in the Milky Way. 44Kb 455 stars

III/180/

N-T-f-v Stephenson C.B., Sanduleak N., Astrophys. J. Suppl. Ser. 33, 459 (1977)

Id.: SS, LS, DM, other

α, δ (1900, 1950) Magnitude, Spectrum Notes and Remarks

An Atlas of low-resolution near-infrared spectra of normal stars 589Kb 61 spectra

III/181/

N-T-f-v Torres-Dodgen A.V., Weaver W.B., Publ. Astron. Soc. Pacific 105, 693 (1993)

Id.: HR or HD

λ, log F_{λ} for 512 points in range 5750-8950Å by exchange from NASA-Goddard

The HDE Charts: accurate positions, proper motions, magnitudes and spectral types of 86933 stars 5.3Mb 86933 stars

III/182/

M-T-f-v Nesterov V.V., Kuzmin A.V., Ashimbaeva N.T., Volchkov A.A., Röser S., Bastian U., Astron. Astrophys. Suppl. Ser. 110, 367 (1995) CANNON A.J., MAYALL M.W., Ann. Harv. Coll. Obs. 112 (1949)

Id.: HD

α, δ (J2000), source

μ_{α}, μ_{δ}, Epoch Magnitude, Sp Remarks See also III/138/

Catalogue Bibliographique de Vitesses Radiales Stellaires 5.2Mb 43821 entries

III/184/

N-T-f-v BARBIER-BROSSAT M., PETIT M., FIGON P., Astron. Astrophys. Suppl. Ser. 108, 603 (1994)

Id.: HD, DM, GCRV, other α, δ (1950) m_{V} , Sp, V_{r}, quality N(measures), dispersion Remarks and References

Supersesdes III/160A/

A Search for Jupiter-Mass Companions to Near-By Stars 39Kb 1082 velocities

III/185/

f-v Walker G.A.H., Walker A.R., Irwin A.W., Larson A.M., Yang S.L.S., Richardson D.C. , Icarus 116, 359 (1995)

Id.: HR, HD, Name

JD, V_{r}, σ (Accuracy 0.1m/s)

Intrinsic energy distribution in stars at 320-760 nm 33Kb 41 sp.classes

III/186/

T-f-v Knyazeva L.N., Kharitonov A.V., Astron. Zh.73, 571 (1996)

Intrinsic energy distribution for 41 spectral types in range B5-M5 by exchange from the Moscow Data Center

IUE Atlas of B-Type Stellar Spectra 5.5Mb 86 stars

III/188/

N-T-f-v Walborn N.R., Parker J.W., Nichols J.S., NASA Reference Publication 1363 (1995)

Id.: HD/HDE, Flamsteed, other Sp SWP, investigator fluxes, splices by exchange from NASA-Goddard See also III/115

Catalogue  References  Contents  Remarks

WEB (Wilson Evans Batten) Catalogue (version 6) 2.9Mb 20795 stars

III/190B/

T-f-v Duflot M., Figon P., Meyssonnier N., Astron. Astrophys. Suppl. Ser. 114, 269 (1995)

Id.: WEB, HD, DM, GCRV, CCDM, HIC α, δ (2000) V_{r}, quality, origin Notes

Catalogue of radial velocities of Nearby Stars 72Kb 200 stars

III/191/

T-f-v Tokovinin A.A., in Catalogue of radial velocity measurements, ed. A.A.Tokovinin. Moscow Univ. Press, 4 (1990) Pis’ma v Astronom. Zh. 16, 52 (1990) Astron. Astrophys. 256, 121 (1992)

Id.: GJ < V_{r} >, errors v sin i, errors Individual (JD, V_{r})

Les vitesses radiales du PPO de FEHRENBACH, années 1955-68 221Kb 2699 stars

III/192/

N-f-v Fehrenbach Ch., Duflot M., Genty V., Amieux G., Bull. Inf. CDS 48, 11 (1996)

Id.: HD, BD α, δ (1950) m_{pg}, Sp V_{r}, quality Reference

Chemical Abundances in Late-Type Stars 171Kb 1108 stars

III/193/

T-f-v Thévenin F., Bull. Inf. CDS 49 (unpublished)

Id.: Field or Cluster θ_{eff}, log g, ζ_{turb} Abundances for elements: Li, O, Na, Mg, Al, Si, Ca, Sc, Ti, V, Cr, Mn, Fe, Co, Ni, Sr, Y, Zr, Mo, Ba, La, Ce, Nd Sm, Eu

Digital Atlas of Optical Spectral Classification of OB Stars 897Kb 78 spectra

III/195/

N-T-f-v Walborn N.R., Fitzpatrick E.L., Publ. Astron. Soc. Pacific 102, 379 (1990)

Id.: HD, alternate Sp 78 FITS spectra in range by exchange from NASA-Goddard A few more spectra

A library of near-IR stellar spectra from 1.428 to 2.5 μm 623Kb 84 stars

III/196/

T-f-v

Lancon  A.,  Rocca-

Volmerange B., Astron. Astrophys. Suppl. Ser. 96, 593 (1992), rev. 1996 Id.: HD, other α, δ (2000) H − K index Comments from literature + 95 FITS spectra

IRAS Low Resolution Spectra (LRS) 5.7Mb 5425 spectra

III/197/

T-f-v Joint IRAS Science Working Group, NASA RP-1190 (1987)

Id.: IRAS α, δ (1950) IRAS data Characterization of spectrum Spectrum between 8 and 22 μm See also II/125

Catalogue  References  Contents  Remarks

The Palomar/MSU nearby star spectroscopic survey 967Kb 1971 stars

III/198/

T-f-v Hawley S.L., Gizis J.E., Reid I.N., Astron. J. 110, 1838 (1995) ; 112, 2799 (1996)

Id.: Seq, Gliese, LHS, ... α, δ (2000) M_{V} , r, Sp Indexes for TiO, CaH, CaOH, Hα

Notes  &  Remarks

The Observed Periods of Ap and Bp stars 501Kb 342 stars

III/199A/

T-f-v Catalano F.A., Renson P., Astron. Astrophys. Suppl. Ser. 127, 421 (1998)

Id.: Rns, HD, HR, other Spwith main elements Periods Remarks, References Notes

Supersedes III/152

A catalogue of [Fe/H] determinations, 1996 ed. 574Kb 5946 stars

III/200B/

T-f-v Cayrel de Strobel G., Soubiran C., Friel E.D., Ralite N., François P., Astron. Astrophys. Suppl. Ser. 124, 299 Id.: HD, Cluster, other Sp, UBV RI, M_{v} Θ_{eff}, log g, Std [F e/H], ref.

Supersedes III/165

See also III/221

The Pulkovo Spectrophotometric Catalog 2.1Mb 609 stars

III/201/

T-f-v Alekseeva G.A., Arkharov A.A. , Galkin V.D. , Hagen-Thorn E.I., Nikanorova I.N., Novikov V.V., Novopashenny V.B., Pakhomov V.P., Ruban E.V., Shchegolev D.E. , Baltic Astron. 5, 603 (1996) ; 6, 481 (1997)

Id.: HR, HD α, δ (2000) V, B − V, Sp Notes Absolute flux in range 320-1080nm, step of 2.5nm (25Å) Catalogue: III/201 (2 tables)

Spectrophotometric Catalogue of Stars 790Kb 1146 stars

III/202/

T-f-v Kharitonov A.V.,Tereshchenko V.M.,Knyazeva L.N., Alma-Ata, Nauka, p. 484,(1988)

Id.: Seq, HR, HD V, Sp F (λ) 3225–7575Å, step of 50Å by exchange from the Moscow Data Center

Emission Lines of the Planetary Nebulae 29Kb 38 PNe

III/203/

f-v Kondratyeva L.N. , Astron. Zh.56, 345 (1979); Astrofizika 22, 153 (1985)

Id.: PK, Other Relative intensities of 45 nebular lines

Catalogue of stars in the Northern Milky Way having H − α in emission 1.9Mb 4174 stars

III/205/

T-f-v Kohoutek L., Wehmeyer R., Abh. Hamburger Sternw. 11, Teil 1 + 2 (1997)

Id.: Seq, HBH, LS, DM, other α, δ (2000); ℓ, b Mag (B, V, or other) Sp, H − α class Remarks

Catalogue  References  Contents  Remarks

13th General Catalogue of MK Spectral Classification 1.6Mb 19208 stars

III/206/

T-f-v Buscombe W., Northwestern Univ., Evanston, Illinois ISBN 0-939160-11-3 (1998)

Id.: HD, other α, δ (2000) Sp UBV by exchange from NASA-Goddard

Moscow Spectrophotometric Catalog 196Kb 223 stars

III/207/

T-f-v Glushneva I.N. et al., Soobshch. Shternberga 219, 3 (1980); Astron. Zh. 57, 1003 (1980) ; Trudy Shternberga 52, 182 (1982); 55, 84 (1983); 61, 272 (1989); 62, 119 (1991)

Id.: HD

F (λ), in λ range 5975– 10825Å by exchange from the Moscow Data Center

Sternberg Spectrophotometric Catalog of Stars 682Kb 866 stars

III/208/

T-f-v Glushneva I.N., Doroshenko V.T., Fetisova T.S., Khruzina T.S., Kolotilov E.A., Mossakovskaya L.V., Ovchinnikov S.L., Voloshina I.B., Trudy Shternberga 53, 50 (1983); 54, 3 (1984); “Spectrophotometry of Bright Stars”, Moscow (1982)

Id.: HD α, δ (2000) UBV, Sp F (λ), in λ range 3225– 7625Å by exchange from the Moscow Data Center

Spectrophotometric Standards of 7-8 mag and Supplement 45Kb 82 stars

III/209/

T-f-v Biryukov V.V., Borisov G.V., Glushneva I.N., Shenavrin V.I., Astron. Astrophys. Transact. 16, N^{∘} 1 (1998); in press (1999)

Id.: HD

α, δ (SIMBAD)

m_{B}, m_{V} , Sp(SIMBAD) F (λ), in λ range 3425– 7525Å by exchange from the Moscow Data Center

Keck/HIRES Sky Line Atlas 783Kb 969 OH-lines

III/211/

T-f-v Osterbrock D.E.,et al., Publ. Astron. Soc. Pacific 108, 277 (1996); 109, 614 (1997)

Molecule, State

λ figures

Catalogue Général de Vitesses Radiales Moyennes pour les étoiles galactiques (Averaged Stellar Radial Velocities for galactic stars) 5.6Mb 36145 stars

III/213/

T-f-v Barbier-Brossat M., Figon P., Astron. Astrophys. Suppl. Ser. 142, 217 (2000)

Id. : HD or BD/CP/CP, other α, δ (1950) m_{V} or m_{B}, Sp \overline{V_{r}}, σ(\overline{V_{r}}), N_{mes} References

Supersedes III/161

Michigan Catalogue of Two-Dimensional Spectral Types for the HD stars, Vol. 5 4.3Mb 31075 stars

III/214/

T-f-v Houk N., Swift C., Department of Astronomy, University of Michigan, Ann Arbor Michigan (1999)

Id.: HD, DM, SAO

Sp, m_{pg}

α, δ (1900); ℓ, b

Notes Zone −12^{∘} < δ < +05^{∘} See also III/31/, III/51/, III/80/, III/133

Catalogue  References  Contents  Remarks

The Seventh Catalogue of Galactic Wolf-Rayet stars 293Kb 226 stars

III/215/

T-f-v van der Hucht K.A. , New Astron. Rev. 45, 135 (2001)

Id.: WR, other α, δ (J2000); ℓ, b Sp; m, color index Period, Binarity Distances, M_{v} References & notes

Supersedes III/85

Atlas of 837.5-877nm spectral region 934Kb 130 spectra

III/217/

T-f-v Andrillat Y., Jaschek C., Jaschek M., Astron. Astrophys. Suppl. Ser. 112, 475 (1995) Carquillat J.M., Jaschek C., Jaschek M., Ginestet N., Astron. Astrophys. Suppl. Ser. 123, 5 (1997)

Id.: HD, HR, Name 1Å-resolution spectra, in range 8375–8770Å

Supersedes III/183

ELODIE archive 407Mb 908 spectra

III/218/

T-f-v Prugniel P., Soubiran, C., Astron. Astrophys. 369, 1048 (2001)

Id.: HD, BD α, δ (J2000) Sp, V_{r} T_{e}ff, logg, [Fe/H] 21 indexes Low-(R=10000) and high- (R=42000)dispersion spectra

A Spectral Library of Galaxies, Star Clusters and Stars 6.5Mb 869 spectra

III/219/

T-f-v Santos Jr. J.F.C., Alloin D., Bica E., Bonatto C. , IAU Symp. 207 “Extragalactic Star Clusters”, held March 12-16 2001, Universidad de Conceptión, Chile Spectra in UV, Optical and near-IR of stars, clusters (galactic, or LMC/SMC, M31) and 240 galaxies Template spectra of stars, clusters, galaxies of various types / morphologies

Catalogue of [Fe/H] 971Kb 6354 stars

III/221A/

T-f-v Cayrel de Strobel G., Soubiran C., Ralite N. , Astron. Astrophys. 373, 159 (2001)

Id.: HD, HIP, Cluster, other α, δ (J2000) Sp, UBV RI, M_{v} T_{eff}, log g, Std [F e/H], ref.

See also III/200

14th General Catalogue of MK Spectral Classification 1.2Mb 17544 stars

III/222/

N-T-f-v Buscombe W. , Northwestern Univ., Evanston, Illinois ISBN 0-939160-12-9 (1999)

Id.: HD/BD/other α, δ (J2000) Sp UBV by exchange from NASA-Goddard Supersedes III/206

MK Spectral Classifications - Twelfth General Catalog 1.3Mb 17438 classifications

III/223/

f-v Buscombe W., Foster B.E. , Northwestern University, Evanston Illinois (1995)

Id.: HD, other α, δ (J2000) Sp, UBV Notes and Remarks

Catalogue  References  Contents  Remarks

Catalogue of Stellar Projected Rotational Velocities 2.2Mb 12000 stars

III/226/

f-v Glebocki R., Gnacinski P., Stawikowski A. , Acta Astron. 50, 509 (2000)

Id.: HD, HR, DM, HIP

α, δ (J2000) m_{v}, Sp v sin i, σ Method used References

Supersedes III/63

Catalogue of Galactic Carbon Stars, 3rd Ed. 1.1Mb 6891 stars

III/227/

f-v Alksnis A., Balklavs A., Dzervitis U., Eglitis I., Paupers O., Pundure I. , Baltic Astron. 10, 1 (2001)

Id.: CGCS (2nd Ed.), names in other lists

α, δ (J2000); ℓ, b BV I (photographic) Sp, Notes and Remarks

Supersedes III/156

Catalog of Cepheid Radial Velocities in 1987-1998 237Kb 126 Cepheids

III/229/

f-v Gorynya N.A., Samus N.N., Sachkov M.E., Rastorguev A.S., Glushkova E.V., Antipin S.V. , Pis’ma Astron. Zh.18, 777 (1992); 22, 198; 24, 939 Id.: Name α, δ (J2000); BV; Sp JD, V_{r}

Catalogue of BV magnitudes and spectral classes for 6000 stars 418Kb 6096 stars

III/230/

f-v Voroshilov V.I., Guseva N.G., Kalandadze N.B., Kolesnik L.N., Kuznetsov V.I., Metreveli M.D., Shapovalov A.N., Ukrainian Acad. Nauk, Kiev, 1985 Id.: Seq α, δ (J2000) (from GSC) BV, Sp

Positions by B. Skiff

The Tycho-2 Spectral Type Catalog 47Mb 351863 stars

III/231/

f-v Wright C.O., Egan M.P., Kraemer K.E., Price S.D. , Astron. Astrophys. 125, 359 (2003)

Id.: TYC, other α, δ (ICRS) Sp, T_{eff} B_{T} , V_{T} Remarks

STELIB: A library of stellar spectra at R about 2000 47Mb 254 stars

III/232/

f-v Le Borgne J.-F., Bruzual G., Pello R., Lançon A., Rocca-Volmerange B., Sanahuja B., Schaerer D., Soubiran C., Vilchez-Gomez R., Astron. Astrophys. 402, 433 (2003)

Id.: HD α, δ (J2000) Observed spectra Corrected spectra (from interstellar absoption and radial velocity)

A Catalogue of Spectroscopically Identified White Dwarfs (Ed. June 2003) 2.1Mb 10132 stars

III/235A/

f-v McCook G.P., Sion E.M. , Astrophys. J. Suppl. Ser. 121, 1 (1999)

Id.: WD, other α, δ (1950) ; mu UBV, uvby, M_{V} Sp, V_{r}, π Notes Supersedes III/210

Catalogue  References  Contents  Remarks

The Byurakan-IRAS Stars (BIS) catalog 165Kb 276 stars

III/237A/

f-v Mickaelian A.M., Gigoyan K.S., Astron. Astrophys. 455, 765 (2006)

Id.: BIS, IRAS, Simbad α, δ (J2000) Sp, R_{pg} BRIJHK mag. IRAS data IRAS data δ > 61^{∘}, |b| > 15^{∘} Data completed in 2006

Synthetic spectra in the near-IR 8500-8750Å 34Mb 952 spectra

III/238/

f-v Munari U., Castelli F., Astron. Astrophys. Suppl. Ser. 141, 141 (2000); Astron. Astrophys. 366, 1003 (2001)

Spectra with a resolution λ/Δλ = 2 × 10^{4} for: −2.5 ≤ log(Z/Z_{⊙}) ≤ +0.5 1.0 ≤ log(g) ≤ 4.5 3500 ≤ T_{eff} ≤ 50000

ISO Astrophysical Spectroscopic Database 2.3Mb 3015 observations

III/242/

f-v Marie Jourdain de Muizon , (in preparation)

Id.: IRAS, ISO-TDT, Simbad names α, δ (J2000) Observation date/time Observing mode λ_{obs}, λ_{lab} Element, transition Comments References

Catalogue of neutral He lines in the spectra of stars. I. 151Kb 546 stars

III/243/

f-v Klochkova V.G., Nasonov D.S., Panchuk V.E. , Special Astrophysical Observatory, Russia (2005)

Id.: HD, BD/CPD Sp W(HeI) λ 4009, 4026, 4121, 4144, 4387, 4438, 4471, 4713

Catalogue of hydrogen line profiles of 235 B, A, F-stars’ spectra 97Kb 235 stars

III/245/

f-v Klochkova V.G., Nasonov D.S. , Special Astrophysical Observatory, Russia (2006)

Id.: HD/BD Obs. Date Equivalent widths and profiles of Hγ and Hδ lines

The first Byurakan spectral sky survey, stars of later spectral classes 374Kb 995

III/246/

f-v Abrahamyan G.V., Gigoyan K.S., Astrofizika 31, 601 (1989); 32, 501 (1990); 33, 317 (1990); 36, 431 (1993); Astrophysics 37, 322 (1994); 38, 115 (1995); 38, 195 (1995); 39, 310 (1996); 40, 131 (1997); 41, 356 (1998); 44, 328 (2001); 45, 322 (2002); 46, 475 (2003)

Id.: FBS L, GSC, IRAS, 2MASS, USNO-B1, other α, δ (J2000) V, R, B, J, H, K, mIRAS, Sp Remarks

Bibliographic catalogue of stellar radial velocities (1991-2006.5). 10Mb 82275 velocities

III/249/

f-v Malaroda S., Levato H., Galliani S. , Complejo Astronómico El Leoncito (CASLEO), San Juan, Argentina (2006)

Id.: HD, CD, CPD α, δ (J2000) V, Sp V_{r}, References Supersedes III/216

Catalogue  References  Contents  Remarks

VVDS-DEEP catalog V1.0 2.7Mb records

III/250/

f-v Le Fèvre O. et al. , Astron. Astrophys. 439, 845 (2005)

Id.: Obs#, IAU α, δ (J2000) z + flags UBV RIJK Links to spectra

Pulkovo Compilation of Radial Velocities for 35495 Hipparcos Stars in a Common System. 1.8Mb 35493 stars

III/252/

f-v Gontcharov G.A. , Astron. Lett. 32, 759 (2006)

Id.: HIP, HD α, δ (J2000) V_{r} + σ N_{obs}

Two-dimentional spectral classification of stars in the direction of the open stellar cluster NGC 2175 1.4Mb records

III/253/

f-v Hovhannesian R.Kh., Epremian R.A., Metrevelii M.D. , Afz (in press)

Id.: Seq, other Sp, Sp(literature)

Astrophysical supplements to the ASCC-2.5. Ia. Radial velocities of 55000 stars and mean radial velocities of 516 Galactic open clusters and associations. 11Mb 54907 stars

III/254/

f-v Kharchenko N.V., Scholz R.- D., Piskunov A.E., Roeser S., Schilbach E., Astron. Nachr. 328, 889 (2007)

Id.: HIP, HD, ASCC-2.5 α, δ (J2000); μ_{α}, μ_{δ} ; V_{r} BV + σ ; Sp Variability, duplicity Quality flags RV Standard Candidates Supersedes III/239

IV. Cross-Identifications

Catalogue  References  Contents  Remarks

A COMPILA-TION OF TRANSIT TABLES FOR STAR NUMBERING IN OPEN CLUSTERS 527Kb 50 clusters

IV/1/

I-M-N-T-f MERMILLIOD J.C.: 1979 , Astron. Astrophys. Suppl. Ser. 36, 163 New Version: 1982 Cross-identifications between different numbering systems for 50 open clusters one open cluster per file.

TABLE OF CORRESPON-DENCES BD/CD/CPD 1.4Mb 48403 records

IV/5/

B-I-M-N-T JUNG J., BISCHOFF M.: 1971, Bull. Inform. CDS 2 Id.: BD, CD, CPD The table has been largely extended since 1971; it is complete for the CSI stars between −40^{∘} and −52^{∘}

Table of Correspondences CSI/ADS/IDS 1.6Mb 45135 records

IV/6/

B-I-M-N-T Jung J., Bischoff M., Ochsenbein F., Bull. Inform. CDS 4, 27 (1973)

Id.: CSI, ADS

α, δ (1900) from IDS CSI is key of catalogue IV/9/ IDS coordinates are not always given when the star belongs to ADS.

TABLE OF CORRESPON-DENCES GICLAS/BD 34Kb 1727 records

IV/7/

B-I-M-N-T MERMILLIOD J.C., Institut d’Astronomie de l’Université de Lausanne & Obs. de Genève (unpublished)

Id.: G, DM

G number is from Lowell Proper Motion Survey, (catalogue I/79/)

Northern sky only

IDENTIFICATION OF HDE STARS 3.2Mb 53521 stars

IV/8/

B-I-M-N-T

( microfiches )

BONNET R.: 1978, Bull. Inform. CDS 14, 114; 15, 115 Id.: HDE, DM Sp

Catalogue of Stellar Identifications, edition 1979 48Mb 434946 stars

IV/9/

B-I-M-N-T

( microfiches )

Ochsenbein F., Bischoff M., Egret D., Astron. Astrophys. Suppl. Ser. 43, 259 (1981)

Cross-identifications

Transit tables See CDS Inform. Bull. 17, 88 (1979) for a complete description 79 microfiches

TABLE OF CORRESPON-DENCES HD-SAO-GC 815Kb 33342 records

IV/10/

I-M-N-T OCHSENBEIN F.: 1981 , CDS, Strasbourg Id.: HD, SAO, GC The file is an excerpt from the CSI (IV/9)

Catalogue  References  Contents  Remarks

CORRESPONDENCES CD/CPD, ZONES −18^{∘} TO −39^{∘} 4.2Mb 171519 records

IV/11/

I-M-N-T BONNET R., OCHSENBEIN F.: 1981 , Bull. Inform. CDS 21, 51 CPD cross-identified with • BD (−18^{∘} to −22^{∘})

• CD (−22^{∘} to −39^{∘})

SAO-HD-GC-DM CROSS-INDEX, VERSION 1983 9.6Mb 258944 stars

IV/12/

B-I-M-N-T-f-v

Roman  N.G.,  War-

ren Jr. W.H., Schofield Jr. N.J., NSSDC/WDC-A-R&S 91-20 MORIN D.: 1973 , Meudon Observatory Id.: SAO, HD, GC, DM α, δ (2000)

By exchange from NASA-Goddard

HD-Dearborn Cross Index 94Kb 5764 stars

IV/13/

I-N-T-f-v Nagy T.A.: 1983, Astron. Data Center Bull. 1, 226 Id.: HD, DO By exchange from NASA-Goddard

LISTE DES BINAIRES SPECTROSCOPIQUES DU FICHIER GENERAL DE TOULOUSE 55Kb 1207 records

IV/16/

B-I-M-N-T-f PEDOUSSAUT A., CAPDEVILLE A., GINESTET N., CARQUILLAT J.M.: 1985, Observatoire de Toulouse Id.: HD, BD, Bayer, Flamsteed, DAO (6^{th} + 7^{th}) , Numbers of Lick Obs. Bulletin Numbers from the Catalogue Complémentaire de Toulouse Prepared at the CDS Supersedes III/65/

Catalogue of HD, HDE, & DM Identifications for stars in open clusters 1.2Mb 345 clusters

IV/17A/

N-T-f-v Mermilliod J.C., Astron. Astrophys. Suppl. Ser. 63, 293 (1976)

Id.: Geneva System, HR, HD, DM, NLS, LSS, GCVS, IDS, ADS, SAO, other

Supersedes IV/2/

WDS-DM-HD-ADS cross index 7.9Mb65467 + 224512 records

IV/18/

N-T-f-v Roman N.G.: 1987, NSSDC/WDC-A-R&S 87-12 Id.: WDS, DM, HD, ADS α, δ (1900)

By exchange from NASA-Goddard

Identification Catalogue of Metal Deficient F-M stars classified spectroscopically. 423Kb 1555 stars

IV/20/

T-f-v Bartkevicius A., Bartkeviciene V.-D., Jodinskiene E., Bull. Vilnius Obs. 76 (1986)

Id.: MDSP/S1, HD, BD/CD, G, LTT other names

α, δ (1950)

By exchange from Moscow Data Center See also III/125

Catalogue  References  Contents  Remarks

INFRARED SOURCE CROSS INDEX 3.1Mb 41009 records

IV/21/

M-N-T SCHMITZ M., MEAD J.M., GEZARI D.Y.: 1987, NASA Ref. Publ. 1182 Id. : source name, IRAS, IRC, AFGL, other α, δ (1950) by exchange from NASA-Goddard

FK5 - SAO - HD - Common Name Cross Index 157Kb 1535 records

IV/22/

f-v Smith W.B. , unpublished (1996)

Id.: FK5, DM, HD, SAO, other α, δ (B1950) μ_{α}, μ_{δ} m_{V} , Sp

Coordinates  and  Iden-

tifications for Sonneberg variables 219Kb 3308

IV/23/

f-v Kinnunen T., Skiff B., Inf. Bull. Var. Stars 4861, 4719, 4720, 4721, 4861, 4862, 4863, 4864, 4865, 4868, 4878, 4879, 4895, 4896, 4897, 4905, 4906 Id.: Sonneberg, GCVS, GSC, IRAS α, δ (J2000)

Catalogue of Galactic Planetary Nebulae 296Mb 1759 PK names

IV/24/

f-v Kohoutek L. , Abh. Hamburger Stermw. XII Id.: PK, Other α, δ (B1950) Discoverers, References Finding Charts Accurate positions

HD identifications for Tycho-2 stars 12Mb 353527 x-ids

IV/25/

f-v Fabricius C., Makarov V.V., Knude J., Wycoff G.L., Astron. Astrophys. 386, 709 (2002)

Id.: HD, Tycho-2 Sp(HD) Duplicity notes Remarks

Der grosse Sternhaufen im Hercules Messier 13 153Kb 1118 stars

IV/26/

f-v Ludendorff H., Publ. Astrophys. Obs. Potsdam, 50 Id.: Ludendorff, Scheiner Δα, Δδ (B1900) m_{pg} Comparison with Scheiner

HD-DM-GC-HR-HIP-Bayer-Flamsteed Cross Index 460Kb 3690 stars

IV/27/

f-v Kostjuk N.D. , Institute of Astronomy of Russian Academy of Sciences (2002)

Id.: ID, DM, GC, HR, HIP, Bayer/Flamsteed α, δ (J2000) Notes

Catalogue  References  Contents  Remarks

Identification of Dahlmark variables 24Kb 213 variables

IV/29/

f-v Skiff B.A., Williams G.V. , Inf. Bull. Variable Stars 4448, 4449, 4450, 4451 (1997)

Id.: LD, IRAS, GSC

α, δ (J2000)

Notes

Cross-identifications of stars with high-proper motions - version 2. 32Mb 488094 stars

IV/30/

f Ivanov G.A. , Kinematika Fiz. Nebesn. Tel., 24, 165 (2008)

Id.: FONAK, AC2000, UCAC2, GSC α, δ (J2000) associated to I/306

V. Combined Data

Catalogue  References  Contents  Remarks

DATA FOR FK4 & SUPPL. STARS 487Kb 3522 records

V/3/

I-M-N-T MORIN D.: 1973, Bull. Inform. CDS 4, 4 Id.: FK4, GC, HD α, δ (1950) ; ℓ, b mu_{α}, mu_{δ}, σ(μ) ; mu_{ℓ}, μ_{b} m, Sp, π, V_{r}

KINEMATIC STUDIES OF EARLY-TYPE STARS 181Kb 1440 + 898 records

V/6/

I-M-N-T RUBIN V.C., BURLEY J.: 1962 , Astron. J. 67, 491 Id.: Number, RCGV, NGC/IC, other m_{v}, V_{r}, M_{v}, A_{v}, Mod. α, δ (1950) ; ℓ, b μ_{α}, μ_{δ} ; μ_{ℓ}, μ_{b} r, R, U, V, W, Z

See also V/31/

SPACE VELOCITIES OF G AND K GIANTS 65Kb 792 records

V/7/

T-M-N YOSS K.M., LUTZ T.E.: 1971 , Mem. Roy. Astron. Soc. 75, 21 Id.: HD/BD M_{V} CN-index, CN-anomaly UV W and ||V⃗ ||

SPACE VELOCITY CATA-LOGUE 384Kb 3483 records

V/8/

I-M-N-T-f-v EGGEN O.J, Roy. Obs. Bull. 51, 79 (1962)

Id.: HD, GC μ_{α}, μ_{δ} ; Sp, V_{r} UV W velocities DU, DV, DW

NEW KINE-MATIC DATA FOR BRIGHT SOUTHERN OB STARS. 57Kb 440 records

V/9/

I-M-N-T LESH J.R.: 1972, Astron. Astrophys. Suppl. Ser. 5, 129 (Table 2)

Id.: HD ; α, δ (1950) ; V, B − V, V_{0} − M_{v} ; r(pc) ; μ_{α}, μ_{δ} ; V_{r}

NEW KINEMATIC DATA FOR BRIGHT SOUTHERN OB STARS: PROPER MO-TIONS 17Kb 440 records

V/10/

I-M-N-T LESH J.R.: 1972, Astron. Astrophys. Suppl. Ser. 5, 129 (Table 4)

{Id.: HD \atop μ_{α}, μ_{δ}}

CATALOGUE OF HIGH VE-LOCITY STARS 78Kb 656 stars

V/11/

I-M-N-T-f EGGEN O.J.: 1964, 65, Roy. Obs. Bulletin Ser. E , 111 Id.: HE, other α, δ (1950) m_{pg}, B − V, U − B, Sp μ_{α}, μ_{δ} V_{r}; UV W velocities

Catalogue  References  Contents  Remarks

Probable members of the Small Magellanic Cloud (New version) 43Kb 524 + 124 stars

V/13/

I-M-N-T-f-v Azzopardi M., Vigneau J. , Astron. Astrophys. Suppl. Ser. 22, 285 (1975)

Id.: Number, chart α, δ (1975) UBV, Sp N(meas.) Known membership The remarks are not available Updated 1978-Oct

uvby estimated astrophysical parameters 1.2Mb 9604 records

V/14/

B-I-M-N-T-f-v

( microfiches )

Philip A.G.D., Egret D. , Astron. Astrophys. Suppl. Ser. 40, 199 (1980)

Id.: Geneva system uvby dereddened colors Sp, M_{V} Direction, log g, [Fe/H]

SAO and Supplementary Data

33Mb 258968 stars

V/15/

T-f-v

( microfiches ) [Data in Simbad]

Ochsenbein F., Bull. Inform. CDS 19, 74 (1980)

Id.: DM, HD, SAO α, δ(1950) μ_{α}, μ_{δ} UBV, Sp, V_{r} Remarks See Bull. Inform. CDS 17, 88 (1979) for a description. Available on microfiche (28mf)

Supergiant Stars 591Kb 5073 records

V/17A/

M-T-f-v

( microfiches )

Egret D., Bull. Inform. CDS 18, 82 (1980)

Id.: CSI, HD α, δ (1950) ; ℓ, b UBV photometry V_{r}, Sp

Catalog of Late-Type Stars with OH, H_{2}O OR SiO Maser Emission 101Kb 351 records

V/18/

N-T-f-v ENGELS D.: 1979, Astron. Astrophys. Suppl. Ser. 36, 337 Id.: HD, other Optical, radio and infrared properties

Catalog of Masses and Ages of stars in 68 open clusters 678Kb 7265 stars

V/19/

T-f-v Piskunov A.E., Bull. Inform. CDS 19, 67 (1980)

Id.: Number in cluster V, B − V, Lum. class log T_{eff}, log L/L_{⊙} Age, Mass By exchange from Moscow Data Center

STELLAR CAT-ALOG FOR ATTITUDE DE-TERMINATIONS IN SPACE. VERSION II.1 10Mb 43099 records

V/21/

I-M-N-T STEIN W.L.: 1980, .

By exchange from NASA-Goddard

Catalogue  References  Contents  Remarks

A CATA-LOGUE OF FOUR-COLOUR PHOTOMETRIC BOXES . 9313 + 11036 boxes

V/23/

M-N-T

( microfiche )

PHILIP A.G.D., EGRET D.: 1983 , Astron. Astrophys. 123, 39 Id.: Geneva system uvbyβ Dereddened colors (m), (c) Sp, Box numbers ONLY on microfiche

CATALOGUE OF GENEVA PHOTOMETRIC BOXES 7.8Mb 79767 records

V/24/

I-M-N-T NICOLET B.: 1982, Astron. Astrophys. Suppl. Ser. 48, 485 For each box: Id.: HD or DM or other α, δ (1900) of central stars α of non-central stars

CSI CATALOGUE WITH SE-LECTED DATA 57Mb 434023 stars

V/26/

I-L-M-N-T-f

[Data in Simbad]

OCHSENBEIN F.: 1983, announced in Bull. Inf. CDS 24, 109

Id.: DM ; α, δ (1950) ; ℓ, b ; μ_{α}, μ_{δ}, V_{r} ; UBV, Sp ; M_{v}, A_{v}

Proper Motions and UBV Photometry of stars in the region of h and χ Per Cluster 286Kb 5504 stars

V/27/

T-f-v Muminov M., Bull. Inform. CDS 24, 95 (1983)

Id.: Number x y μ_{x} μ_{y} V, B − V, U − B Membership probability By exchange from Moscow Data Center Clusters, open; Proper motions; Photometry, UBV; Associations, stellar; Regional catalog

PHOTOMETRY AND PROPER MO-TIONS OF STARS IN FOUR OPEN CLUSTERS IN CAS-SIOPEA 130Kb 2160 records

V/28/

N-T-f-v FROLOV V.N.: 1982, Pulkovo Obs.

Id.: Number x y μ_{x} μ_{y} V, B − V, U − B Cluster membership Region of NGC 7788, NGC 7790, Berkeley 58 and anonymous

Catalog of space velocity vectors of 2595 B and A stars 177Kb 2595 stars

V/30/

I-M-N-T-f-v Palous J., Bull. Inform. CDS 23, 96 (1982) Bull. Astron. Inst. Czech. 34, 286 Id.: number ℓ, b, distance UV W, m_{v}, Sp

Catalogue of Kinematic data for O-B5 Stars 227Kb 1440 records

V/31A/

M-N-T-f-v Rubin V.C., Burley J., Kiasatpoor A., Klock B., Pease G., Rutscheidt E., Smith C., Astron. J. 67, 491 (1962)

Id.: Wilson, HD, DM, other α,δ (1950) ; ℓ, b μ_{α}, μ_{δ} ; μ_{ℓ}, μ_{b} m_{v}, V_{r}, M_{v}, A_{v}, m − M r, R, U, V, W, z References Modified version of V/6/

Catalogue  References  Contents  Remarks

Catalogue of stars within 25 parsecs of the Sun 321Kb 2150 stars

V/32A/

N-T-f-v

[Identifiers in Simbad]

Woolley R., Epps E.A., Penston M.J., Pocock S.B. , Royal Obs. Annals 5 (1970)

Id.: Number, Yale, HD, DM, BS, Vyss, Wilson α, δ (1900, 1950) UBV ; M_{v}, Sp μ_{α}, μ_{δ}, V_{r}, π UV W velocities Orbit parameters ϖei Comments

DDO Photometry and Radial Velocities of Stars at Galactic Poles 281Kb 4079 records

V/33/

I-M-N-T-f-v Hartkopf W.I., Yoss K.M. , Astron. J. 87, 1679 (1982)

Id.: HD, other DDO and Mg photometry V_{r}

Supplement to the Bright Star Catalogue 708Kb 2603 stars

V/36B/

N-T-f-v Hoffleit D., Saladyga M., Wlasuk P., Yale University Observatory, New Haven, Connecticut (1983)

Id.: HD, SAO, ADS α, δ(1900, 2000) ; ell, b UBV RI, Sp μ_{α}, μ_{δ}, π, V_{r}, v sin i Binarity, Variability Notes From NASA / Goddard Stars with V <= 7.10 not already in the BSCat

CATALOGUE OF STELLAR UV FLUXES 8.2Mb 31215 records

V/38/

I-M-N-T THOMPSON G.I., NANDY K., JAMAR C., MON-FILS A., HOUZIAUX L., CARNOCHAN A., WIL-SON R.: 1979, Bull. Inform. CDS 17, 78 Id.: S2/68, HD, DM, SAO, HR, Name α, δ, μ_{α}, μ_{δ} m_{v}, UBV, uvbyβ, Sp V_{r}, v sin i F (λ) for λ (width 330Å): 1565, 1965, 2365, 2740, 2740 Å σ(F_{λ})

By exchange from NASA-Goddard See also II/59

The Fourth Catalog of Orbits of Visual Binary Stars 292Kb 928 orbits

V/39/

N-T-f-v Worley C.E., Heintz W.D., Publ. U.S. Naval Obs. (2) 24, part VII (1983)

Id.: Name, ADS α, δ (1900) m_{v}(A), m_{v}(B) ; Sp(A), Sp(B) Orbit parameters References & Notes By exchange from NASA-Goddard

CATALOGUE OF PHYSICAL PARAME-TERS OF SPECTROSCOPIC BINARY STARS 171Kb 936 binaries

V/41/

I-M-N-T-f KRAICHEVA Z., POPOVA E., TUTUKOV A., YUN-GELSON L.: 1980, Bull. Inform. CDS 19, 71 Id.: HD, BD, Name α, δ(1900) V, Sp Q, P, Orbit parameters K, V_{0}, f(m), log M_{1}, log M_{2}, log a, log J, log j By exchange from Moscow Data Center Previously under number III/57/

Catalog of orbital elements, masses and luminosities of close double stars 778Kb 11725 records

V/42/

I-M-N-T-f-v Svechnikov M.A., Bessonova L.A., Bull. Inform. CDS 26, 99 (1984)

Id.: Variable name Photometric elements of EBs Absolute elements of EBs Notes & references By exchange from Moscow Data Center Previously under number III/82/

Catalogue  References  Contents  Remarks

CATALOGUE OF MASSES AND AGES OF STARS IN TWELVE OPEN CLUSTERS 129Kb 1027 stars

V/43/

I-M-N-T-f MYAKUTIN V.I., RAM SAGAR JOSHI U.C.: 1984, Bull. Inform. CDS 26, 103 Id.: Number V, B − V, U − B, E(B − V) Membership probability Luminosity class log T_{eff} , log L/L_{⊙} Age, M/M_{⊙} By exchange from Moscow Data Center Previously under number VII/45/

CATALOGUE OF STARS IN STELLAR ASSOCIATIONS AND YOUNG CLUSTERS 608Kb 2322 stars

V/44/

I-L-M-N-T-f HUMPHREYS R.M., McELROY D.B.: 1984, Univ. of Minnesota Id.: HD, other Sp, ℓ, b V, B − V, U − B, E(B − V) M_{v}, M_{bol}, log T_{eff} Association, comments Previously under number VII/46

A CATALOGUE OF FIELD TYPE II CEPHEIDS 15Kb 208 records

V/48/

I-M-N-T HARRIS H.C.: 1985, Astron. J. 90, 756 Id.: Variable name α, δ ; ℓ, b m, P, distance r, z E(B − V)

By exchange from NASA-Goddard

14ème CATALOGUE COM-PLEMEN-TAIRE DE BINAIRES SPEC-TROSCOPIQUES 115Kb 384 binaries

V/49/

I-M-N-T PEDOUSSAUT A., GINESTET N., CARQUIL-LAT J.M.: 1984, Astron. Astrophys. Suppl. Ser. 58, 601 Id.: Number, HD, BD, Batten α, δ (1900) m, Sp Orbital parameters Keypunched at the CDS see also V/60/

The Bright Star Catalogue, 5th Revised Ed. (Preliminary Version) 2.2Mb 9110 stars

V/50/

N-T-f-v

[Identifiers in Simbad]

Hoffleit D., Warren Jr W.H.: 1991, Id.: HR, HD, DM, ADS, Flamsteed, Bayer α, δ (1900, 2000) ; ℓ, b UV BRI photometry μ_{α}, μ_{δ}, π Sp, V_{r}, v sin i Duplicity, Variability Notes & Remarks Supersedes V/25 by exchange from NASA-Goddard

Data Inventory of space-based Celestial Observations (DISCO 1.1) 8.7Mb 74952 records

V/51/

I-N-T-f-v BROTZMAN L.E., HILL R.S., MEAD J.M.: 1986, Astron. Data Center Bull. 1 α, δ Mission or Catalog References By exchange from NASA-Goddard

Photometry and Proper Motions of stars in the open cluster M67 141Kb 1068 stars

V/52/

I-M-N-T-f-v Frolov V.N., Ananyevskaya J.K.: 1986 , Moscow Id.: Number, Sanders, Fagerholm x y, μ_{x} μ_{y} UBV (V ≤ 14.5) Membership probability By exchange from Moscow Data Center See also II/123/

Catalogue  References  Contents  Remarks

Catalogue of the Brightest Stars 438Kb 1628 stars

V/53A/

N-T-f-v Ochsenbein F., Halbwachs J.L., Bull. Inform. CDS 32, 83 (1987) Ochsenbein F., Acker A., Legrand E., Poncelet J.M., Thuet-Fleck E., “Le catalogue des étoiles les plus brillantes”, A.Acker (ed.), Strasbourg Obs. (1988)

Id.: HR, HD, HIP, name α, δ (2000s, ICRS), precession λ, β(ecliptic), ℓ, b μ_{α}μ_{δ}, V_{r}, π, Distance UBV RI photometry uvbyβ photometry V_{rot}, H_{γ}, Sp For multiple systems: Δm, ρ, N(components)

Hipparcos

proper motions and parallaxes added

A Catalogue of high-velocity stars 116Kb 633 stars

V/55/

I-M-N-T-f-v Roman N.G., Astrophys. J. Suppl. Ser. 2, 195 (1955) Contrib. McDonald Obs. 262 Id.: HD, BD, RHV, MDSP/S1 αδ(1950, 1900, 1875, 1855)

ell, b; μ, V_{r}; π; UBV π(spectr.), UV W vel. Galactic orbit Notes & references By exchange from Moscow Data Center

The Catalogue of metallicities, velocity components and orbital parameters for F2-K5 dwarfs in the vicinity of the Sun 132Kb 1143 stars

V/58/

I-M-T-f-v Marsakov V.A., Shevelev Yu.G.: , Rostov State Univ., USSR Id.: HD (ell, b), r, UV W UBV, δ(U_{B}) [F e/H], Sp Orbit (a, e) param.

By exchange from Moscow Data Center See also V/89

15ème Catalogue Complémentaire de Binaires Spectroscopiques 117Kb 436 binaries

V/60/

T-f-v Pédoussaut A., Carquillat J.M., Ginestet N., Vigneau J., Astron. Astrophys. Suppl. Ser. 75, 441 (1988) Malkov O.Yu., Bull. Inform. CDS 36 (1989)

Id.: HD, name α, δ (1900) Sp, m_{V} , m_{B} P, T, ω, e, K, V_{0}, a sin i, m sin^{3} i, f(m)

Reference see also V/49/

Ptolemy’s Star Catalogue 54Kb 1027 stars

V/61/

N-T-f-v MANITIUS K.: 1913, B.G. Teubner, Leipzig JASCHEK C.: 1987 , Bull. Inform. CDS 33, 145 Id.: HR, Constellation longitude, latitude, m Disagreement in identification Keypunched at the CDS

POPPER’S CATALOGUE OF ASTROPHYSICAL PARAM-ETERS 62Kb 131 binaries

V/63/

B-I-M-N-T POPPER D.M.: 1980, Ann. Rev. Astron. Astrophys. 18, 115 MALKOV O. Yu.: 1989 , Bull. Inform. CDS 37, 175 Id.: HD, DM, Name P eriod, Sp, M_{v} log T_{eff}, Masses, Radii, log L, σ Remarks & references Keypunched at the CDS

Catalogue  References  Contents  Remarks

A Spectroscopically Selected calalog of field K giants in the Galactic halo 16Kb 106 stars

V/65/

N-T-f-v Ratnatunga K.U., Freeman K.C., Astrophys. J. 291, 260 (1989) Astrophys. J. 339, 126 Id.: Ratnatunga, Eriksson, Bok&Basinski BV magnitudes, M_{V} α, δ (1950) ; (ℓ, b) V_{r}, [Fe/H] by exchange from NASA-Goddard

An astrometric catalogue of radio stars 195Kb 186 radio stars

V/67/

B-M-N-T-f-v Hering R., Walter H.G., de Vegt C., Bull. Inf. CDS 38, 209 (1990)

Id.: HD, BD, SAO, AGK3, other α, δ (optical + radio) μ(optical + radio) π(optical + radio) V_{r}, m, Sp Types of variability Guides to maps Observation priorities The file is in printout format

A Catalog of observations for stellar Masers 1.2Mb 3647 sources

V/68A/

N-T-f-v Benson P.J., Little-Marenin I.R., Attridge J., Blais K., Rudolph D., Woods T., Rubiera M., Keefe H., Astrophys. J. Suppl. Ser. 74, 911 (1990)

Id.: Number, Variable, IRAS, other α, δ (1950) V_{LSR}, Sp Variable type, P H_{2}O, SiO, OH detections The electronic version contains more sources than the printed one

MK Classification and Photometry of stars used for Time and Latitude observations at Mizusawa and Washington (Revised Version) 109Kb 591 stars

V/69/

N-T-f-v Sato K., Kuji S. , Nat. Astron. Obs. of Japan, Mizusawashi, Iwate-ken, 023 Japan; Astron. Astrophys. Suppl. Ser. 85, 1069 (1990)

Id.: Number, BD, HD, other UBV + Remarks Spectral classes Obs. date Plate Quality by exchange from NAO-Japan

CATALOGUE OF NEARBY STARS, 3rd edition (Preliminary Version) 744Kb 3803 stars

V/70A/

N-T-f-v GLIESE W., JAHREISS H.: 1991 , Astron. Rechen-Institut, Heidelberg Id.: Gliese, GJ, HD, DM, G, LHS, ... α, δ (1950) (μ, θ), V_{r}, Sp, π UBV RI colour indices M_{V} + quality UV W velocities by exchange from NASA-Goddard Supersedes V/35

A PRELIMINARY CAT-ALOGUE OF FIELD POPU-LATION II STARS (POP2) 8.6Mb 34974 stars

V/72/

N-T BARTKEVICIUS A., BARTKEVICIENE D.: 1993 , Bull. Inf. CDS 42, 19 Id.: list in a separate file α, δ (1950) + source ℓ, b Literature by exchange from Moscow Data Center

Catalogue  References  Contents  Remarks

Third Catalogue of emissionline stars of the Orion population 383Kb 742 stars

V/73A/

T-f-v Herbig G.H., Bell K.R., Lick Obs. Bull. 1111 (1988)

Id.: HBC, Var. name, other α, δ (1950) + ref., ℓ, b m, UBV RI, Sp, refs. W (Hα), ref. Location by exchange from NAO-Japan

CATALOGUE OF ASTROPHYSICAL PARAM-ETERS OF BINARY SYS-TEMS 43Kb 144 binaries

V/75/

B-N-T-f MALKOV O.Yu: 1993, Bull. Inf. CDS 42, 27 Id.: HD / Name Period, Sp, M_{v} log T_{eff} with σ M with σ Radius with σ log L with σ log g with σ Remarks & References

A Catalog of Chromospherically Active Binary Stars (2nd Edition) 528Kb 206 binaries

V/76/

N-T-f-v Strassmeier K.G., Hall D.S., Fekel F.C., Scheck M., Astron. Astrophys. Suppl. Ser. 100, 173 (1993)

Id.: Number, Variable, HD, HR, ... α, δ (2000) , XY Z, UV W UBV RI , Sp v. sin i, F (X), F (radio) Orbital Elements M_{v}, r, M, R Notes & References

Supersedes V/71/

CATALOGUE OF THE OB-JECTS IN THE DIRECTION OF THE SMALL MAGEL-LANIC CLOUD 2.6Mb 16370 objects

V/78/

N-T-f BISCHOFF M., FLORSCH A., FLORSCH J., MAR-COUT J.:1988 , Publ. Obs. Astron. Strasbourg, Equipe ”Populations Stellaires”

Id.: HD, GCVS, CPD, CSI,... α, δ (1975) UBV, Sp, V_{r} Ref., Remarks Variables Non-stellar objets ξ, η Bibliography

Membership of Low-Mass Stars in the Open Cluster Alpha Perse 343Kb 327 stars

V/80/

T-f-v Prosser C.F. , Astron. J. 103, 488 (1992)

Id.: HE α, δ (B1950) m_{V} , m_{B} Membership Pr. V_{r} References and notes by exchange from NASA-Goddard

Taxonomy of Barium stars 79Kb 191 stars

V/81/

N-T-f-v Lu P.K., Astron. J. 101, 2229 (1991)

Id.: HD/CD/CoD, SAO α, δ (1900); (ℓ, b Sp, Ba intensity class UBV RI , DDO Colors V_{r}, μ_{α}, μ_{δ} UV W velocities Remarks & Notes by exchange from NASA-Goddard

Catalogue  References  Contents  Remarks

Strasbourg-ESO Catalogue of Galactic Planetary Nebulae 1.1Mb 1143 PNs

V/84/

N-T-f-v Acker A., Ochsenbein F., Stenholm B., Tylenda R., Marcout J., Schohn C., European Southern Observatory ISBN 3-923524-41-2 (1992)

Id.: PN G, PK, IRAS,... α, δ (1950) Diameters (Opt.&Radio) V_{r}; Expansion Velocities F(Hβ); HKL photom. Intensities for 8 lines IRAS fluxes Distance estimations IUE observations list Notes \underline{Central Star} data Lists of Possible & Rejected PNs

Supersedes VII/24/

Stellar Mass Catalogue. Preliminary Version 58Kb 246 systems

V/85A/

N-T-f-v Belikov A.N., Bull. Inf. CDS 47, 9 (1995)

Id.: two identifiers M_{1}, M_{2}, σ(M_{1}), σ(M_{2}), method Sp, System type Notes & References

DISTRIBUTION OF INTER-STELLAR MATTER IN THE GALACTIC DISK FROM VI-SUAL EXTINCTION DATA 4.2Mb 15999 stars

V/86/

M-T-f-v GUARINOS J., Ph.D. Thesis, Strasbourg Observatory in Astronomy from Large Databases II ISBN 3-923524-47-1, 301 Id.: HD or DM Galactic (ℓ, b) Sp+ source UBV, M_{V} (B − V)_{0} , A_{V} r, XY Z distances Stars with peculiar colour excesses

Averaged [Fe/H] for Cool Dwarfs 50Kb 497 stars

V/88/

T-f-v Taylor B.J., Publ. Astron. Soc. Pacific 106, 704 (1994) Publ. Astron. Soc. Pacific 107, 734 (1995)

Id.: BD, HD, Wolley [Fe/H], σ References by exchange from NASA-Goddard

Catalogue of Ages, Metallicities, Orbital elements and other parameters for nearby F stars 791Kb 5489 F stars

V/89/

N-T-f-v MARSAKOV V.A., SHEVELEV Yu.G., Bull. Inf. CDS 47, 13 (1995) Astron. Zhurn. 70, 1218 (1993)

Id.: HD (b − y)_{cor}, δm_{1}, δc_{1} [F e/H], T_{eff}, log g M_{V} , Age, Sp ℓ, b UV W vel. Galactic orbit σ

Catalogue of open cluster parameters from UBV-data 96Kb 340 clusters

V/96/

T-f-v Loktin A.V., Matkin N.V., Gerasimenko T.P., Astron.Astrophys.Trans. 4, 153 (1994)

Id.: Cluster designation Distance, A_{V} , ages, [Fe/H] by exchange from the Moscow Data Center

Catalogue  References  Contents  Remarks

Updated 3rd Cat Variable Stars in Globular Clusters 682Kb 2664 variables

V/97/

T-f-v Sawyer Hogg H., Publ. David Dunlap Observatory, 3, No. 6 (1973) Clement C., Unpublished (1997)

\underline{Clusters}: α, δ (1950) \underline{Variables}: x, y Mag: min, max, epoch, P Remarks by exchange from NASA-Goddard

MSX Infrared Astrometric Catalog 34Mb 177860 sources

V/98/

T-f-v Egan M.P., Price S.D., Astron. J. 112, 2862 (1996)

Id.: (astrometric) (IR) α, δ (J2000) Sp, m_{v}, LRS IRAS fluxes (12,25,60,100 μm) SPIRITIII estimated fluxes

See also V/114

Catalogue of nearest stars until 10pc 50Kb 356 stars

V/101/

T-f-v Zakhozhaj V.A., Vestnik Khar’kovskogo Universiteta 190, 52 (1979)

Id.: Zkh, Gliese, other α, δ (J2000) μ_{α}, μ_{δ}, V_{r}, π m_{v}, M_{v}, Sp Mass, Radius by exchange from the Moscow Data Center ‘

Radial velocities and DDO, BV photometry of HD stars near NGP 279Kb 914 stars

V/103/

T-f-v Yoss K.M., Griffin R.F., J. Astrophys. Astron. 18, 161 (1997)

Id.: HD / BD, Upgren V_{r}, obs. Date Sp, BV, C_{1}C_{2}C_{3} μ_{α}, μ_{δ}, Mv, [Fe/H] Kin. zUV W ae

A  General  Cata-

logue of Herbig-Haro Objects,

2. Edition 938Kb 454 objects

V/104/

T-f-v Reipurth B., Center for Astrophysics, Univ. Colorado, USA (1999)

Id.: HH, other α, δ (1950), Region Distance Extensive Notes

Catalogue of X-ray binaries 604Kb 130 binaries

V/106/

T-f-v Liu Q.Z., van Paradijs J., van den Heuvel E.P.J., Astron. Astrophys. Suppl. Ser. 147, 25 (2000); Astron. Astrophys. 368, 1021 (2001)

Id.: X-ray, optical counterpart α, δ (J2000); ℓ, b F (X), UBV, Sp Periods Notes and references Supersedes V/90 HMXB part: J/A+A/455/1165

SKY2000 Master Catalog, Version 4 115Mb 299167 stars

V/109/

f-v Myers J.R., Sandé C.B., Miller A.C., Warren Jr. W.H., Tracewell D.A., Goddard Space Flight Center, Flight Dynamics Division (1999)

Id.: HD, Skymap, SAO, PPM, WDS, DM, HR, Name α, δ (J2000); ell, b μ, π, V_{r}, m_{pg} UBV, Sp, colours Neighbor star, variability Supersedes V/105/

Catalogue  References  Contents  Remarks

Catalog of Star-Forming Regions in the Galaxy 11Mb 3235 regions

V/112/

f-v Avedisova V.S , Astron. Zh.79, 216 (2002)

Id.: SFR Name, IRAS, other α, δ (radio J2000), ℓ, b Dimensions Radio or IR Flux Names and References

Supersedes V/74

MSX6C Infrared Point Source Catalog 191Mb 529724 sources

V/114/

f-v Egan M.P., Price S.D., Kraemer K.E., Mizuno D.R., Carey S.J., Wright C.O., Engelke C.W., Cohen M., Gugliotti G. M. , Air Force Research Laboratory Technical Report AFRL-VS-TR-2003-1589 Id.: MSX6C α, δ (J2000) IR Fluxes 4–25μm + σ (6 bands B_{1}B_{2}ACDE) S/N, Qualities, variability flags

Supersedes V/107

Parameters of Algol-type eclipsing binaries 82Kb 232 systems

V/115/

f-v Surkova L.P., Svechnikov M.A. ,

Id.: GCVS ; P, i, M_{1}, M_{2}, R_{1}, R_{2} ; M_{bol:1,2}

A galactic O star catalog 233Kb 378 stars

V/116/

f-v Maiz-Apellaniz J., Walborn N.R., Galue H.A., Wei L.H., Astrophys. J. Suppl. Ser. 151, 103 (2004)

Id.: GOS, HD, TYC, WDS, other α, δ (J2000); ℓb UBV JHK, uvby, Sp Notes and References

The Geneva-Copenhagen survey of the Solar neighbourhood 4.7Mb 16682 stars

V/117/

f-v Nordstrom B., Mayor M., Andersen J., Holmberg J., Pont F., Jorgensen B.R., Olsen E.H., Udry S., Mowlavi N. , Astron. Astrophys. 418, 989 (2004)

Id.: HIP, HD/BD/CD/CP

α, δ (J2000); (ℓb) μ_{α}, μ_{δ}, V_{r}; UV W Distance (r_{⊙}, r_{GC}) T_{eff}, [Fe/H], Mass, Age m_{V} , M_{V} , b − y, β Galactic orbit

Catalogue of orbital elements, masses and luminosities of eclipsing binaries with detached main sequence components 22Kb 44 binaries

V/118/

f-v Perevozkina E.L., Svechnikov M.A. , I.I. Bondarenko Ed., 1999, Ekaterinburg, Ural University Id.: Seq, GCVS P, i M_{1}, M_{2}, L_{bol}

Catalogue of field contact binary stars 97Kb 361 binaries

V/119/

f-v Pribulla T., Kreiner J.M., Tremko J. , Contr. Astron. Obs. Skalnate Pleso 33, 38 (2003)

Id.: Variable name α, δ (J2000) P, Class of system V_{min}, V_{max}, Sp Mass ratio Notes and References

Catalogue  References  Contents  Remarks

Catalogue of Photometric, Geometric, and Absolute elements of contact binary stars of the early spectral type 45Kb 148 binaries

V/120/

f-v Bondarenko I.I., Perevozkina E.L. , Publ. Odessa Obs. 9, 20 (1996)

Id.: Variable name Sp of Components Masses, radiuses Roche-lobe filling Notes

Catalog of Detached Main-Sequence Eclipsing Binaries 50Kb 113 systems

V/121/

f-v Svechnikov M.A., Perevozkina E.L. , Ekaterinbourg, Ural University, p. 5 (1999)

Id.: GCVS Period, Sp Masses, Radii M_{bol} References

A catalog and atlas of cataclysmic variables: the Final edition 707Kb 1740 CV*

V/123A/

f-v Downes R.A., Webbink R.F., Shara M.M., Ritter H., Kolb U., Duerbeck H.W. , Publ. Astron. Soc. Pacific 113, 764 (Version 2006 Feb.)

Id.: GCVS, other α, δ (2000) Year of outburst m_{min}, m_{max} References

Supersedes V/110

Catalogue of approximate photometric and absolute elements of eclipsing variable stars 811Kb 3793 binaries

V/124/

f-v Svechnikov M.A., Kuznetsova Eh.F., Ural university publication, Sverdlovsk (1990)

Id.: Variable name α, δ (J2000) P, Sp, orbital elements, masses, radiuses Notes and References

Photometric and Spectroscopic Databases for Luminous Stars

14Mb 18693 stars

V/125/

f-v Reed C. , Astron. J. 130, 1652 (2005)

Id.: LSS, HD, DM, other ℓ, b UBV, Hβ V_{r}, Sp Notes and References

Supersedes V/92

Compiled catalogue of stellar proper motions in the MEGA program areas. 2.1Mb 18169 stars

V/126/

f-v Kharchenko N.V. , Kinematika Fiz. Nebesn. Tel. 20, 366 (2004)

Id.: ASCC-2.5, HIP, HD, DM, MEGA-G α, δ (J2000) μ_{α}μ_{δ} + σ BV, Sp

The Macquarie/AAO/Strasbourg Hα Planetary Nebula Catalogue (MASH) 1027Mb records

V/127/

f-v Parker Q.A., Acker A., Frew D.J., Hartley M., Peyaud A.E.J., Phillipps S., Russeil D., Beaulieu S.F., Cohen M., Koppen J., Marcout J., Miszalski B., Morgan D.H., Morris R.A.H., Ochsenbein F., Pierce M.J., Vaughan A.E., Mon. Not. R. Astron. Soc. 373, 79 (2006)

Id.: MASH, PNG α, δ (J2000) σ

VI. Miscellaneous Data

Catalogue  References  Contents  Remarks

CATALOGUE OF PHOTO-METRIC SEQUENCES 46Kb 412 records

VI/1/

I-M-N-T ARGUE A.N., BOK B.J.: 1973 , Mineograph Arizona University Id.: Number α, δ (1950), ℓ, b Range in V magnitude Photometric system References

SENSITIVITY FUNCTIONS OF PHOTOMETRIC SYS-TEMS 44Kb 30 records

VI/5/

I-M-N-T HAUCK B., MERMILLIOD J.C.: 1976 , Bull. Inform. CDS 10, 28 Tables of sensitivity functions (constant step for each filter)

A list of the systems is given in the reference paper (table 1)

A Table of semiempirical gf values 20Mb 265587 records

VI/10/

I-M-N-T-f-v Kurucz R.L. & Peytremann E., SAO Special Report 362 (1975) λ, log g_{f}, E_{1}, E_{2} Element & configurations

BIBLIOGRAPHICAL INDEX OF PLANETARY NEBULAE FOR THE PERIOD 1965-79 . 1429 records

VI/14/

M-N-T

( microfiches )

ACKER A., MARCOUT J., OCHSENBEIN F.: 1981, Astron. Astrophys. Suppl. Ser. 43, 265 2 files:

• exhaustive title of reference and running number • list of running numbers for each P.N.

Available only on microfiche.

Line spectra of the elements 2.8Mb 46663 lines

VI/16/

N-T-f-v Reader J. & Corliss Ch.H., 61st ed., CRC Handbook of Chemistry & Physics (1980– 81)

Contains the outstanding lines of atoms from vacuum ultra-violet to the far infrared. Last file is a finding list of the table compiled by Didelon P. (CDS)

Tape produced by U.S., National Bureau of Standards

UBV PHOTOELECTRIC SE-QUENCES AND SPECTRAL CLASSIFICATION IN SA92-115 81Kb 642 stars

VI/19/

N-T-f-v LANDOLT A.U.: 1973, Astron. J. 78, 959 DRILLING J.S., LANDOLT A.U.: 1979, Astron. J. 84, 783 Id.: from Harvard Annals 101 α, δ (1975) UBV, σ, notes Sp

Palomar Observatory Sky Survey plate data (POSS-I) 164Kb 1037 plates

VI/25/

N-T-f-v National Geographic Society & Palomar Obs.: 1960, Sky Survey Catalogue of Plates Id.: POSS, MLP α, δ (1950) of plate center limits in α ℓ, b of plate center Number of SAO stars date, exposure, hour angles By exchange from NASA-Goddard

Catalogue  References  Contents  Remarks

A Finding List for the Multiplet Tables of NSRDS-NBS 3, Section 1-10 704Kb 9969 records

VI/29A/

N-T-f-v Adelman C.J., Adelman S.J., Fischel D., Warren W.H. , Astron. Astrophys. Suppl. Ser. 60, 339 (1985)

Element, Atomic number λ, intensity Excitation potentials Multiplet number Reference By exchange from NASA-Goddard

Plate centers of the ESO Sky Survey 157Kb 606 fields

VI/30A/

N-T-f-v Holmberg E.B., Lamberts A., Schuster H.E., West R.M. , Astron. Astrophys. Suppl. Ser. 18, 463 (1974) α, δ (1950) of field centers Observation dates Image qualities By exchange from NASA-Goddard .

PLATE CENTERS OF THE LICK SKY SURVEY 22Kb 212 plates

VI/31/

N-T-f-v LICK OBS.: 1965; DOUGHTY et al.: 1972, Univ. of Canterbury Id.: Field number number of CSI stars α, δ (1950) of the field centers epoch By exchange from NASA-Goddard

The Bidelman-Parsons Spectroscopic and Bibliographical catalog 12Mb 40312 stars

VI/32/

N-T-f-v PARSONS S.B., BUTA N.S., BIDEL-MAN W.P.: 1980, Bull. Inform. CDS 18, 86 Id.: HD, DM α, δ (1950) remarks references

SEMIEMPIRICAL CAL-CULATION OF gf VALUES FOR Fe II 25Mb 431933 records

VI/33/

I-M-N-T-f KURUCZ R.L.: 1981, SAO Special Report 390 λ, log g_{f} energy level configuration By exchange from NASA-Goddard

A Multiplet table for MnI 1.4Mb 8724 lines

VI/34/

I-L-N-T-f-v Adelman S.J., Svatek G.F., van Winkler K., Warren W.H., Astron. Astrophys. Suppl. Ser. 80, 285 (1989) λ, reference code Laboratory intensity Excitation potentials J values, configuration codes, multiplet number By exchange from NASA-Goddard

Annoted Bibliography of Multivariate Statistical Methods in Astronomy 114Kb 2951 records

VI/37/

I-M-N-T-f-v Murtagh F., Heck A., Astron. Astrophys. Suppl. Ser. 68, 113 (1987) Bull. Inform. CDS 31, 183 Introduction 169 references Index of authors 1st file in LaTeX format

MODEL ATMOSPHERE FOR G, F, A, B, AND O STARS 13Mb 187033 records

VI/39/

I-M-N-T-f-v KURUCZ R.L.: 1979, Astrophys. J. Suppl. Ser. 40, 1 For each model: λ (nm) H_{ν} H_{ν}^{cont} See also VI/77/

Catalogue  References  Contents  Remarks

THE RE-VISED YALE ISOCHRONES AND LUMINOSITY FUNC-TIONS 10Mb 24 sets

VI/40/

I-M-N-T-f GREEN E.M., DEMAR-QUE P., KING C.R.: 1987,, Yale University Models for 24 sets (Z, Y) Age, mass, log T_{eff}, log L UBV RI Fortran programs By exchange from NASA-Goddard

Constellation Boundary Data 26Kb 357 records

VI/42/

N-T-f-v Roman N.G., Rappaport B.N. , Publ. Astron. Soc. Pacific 99, 695 (1987)

Boundaries in α_{1875}δ_{1875} Fortran and C program to compute constellation name from position See also VI/49

EXOSAT OBSERVATION LOG 306Kb 2355 records

VI/43/

I-M-N-T-f-v Sternberg J.R., White N.E., Barr P., Osborne L., Exosat Observation log, Noordwijk: ESTEC Printingservice, December 1986 Times of stable pointing α, δ (1950) source name observation mode PI names By exchange from NASA-Goddard Supersedes VI/41/ and VI/36/

Finding list for observers of Interacting Binary Stars (5th Ed.) 1.1Mb 3564 systems

VI/44/

T-f-v Wood F.B., Oliver J.P., Florkowski D.R., Koch R.H., Publ. Dept. of Astron., Univ. of Florida, 1 (1980) Publ. Univ. Pennsylvania Astron. XII Id.: BD, CoD, CpD, HD α, δ(1900) m_{B} at maximum depth of minima duration of minimum Sp, epoch, period By exchange from NASA-Goddard see also VI/23/

A BIBLIOGRAPHICAL CAT-ALOGUE OF RR LYRAE STARS 352Kb 6367 stars

VI/45/

B-I-M-N-T-f HECK A.: 1988, Bull. Inform. CDS 34, 137 Id.: Cross-identifications Bibliographical references

Supersedes VI/8/

Second Catalog of Interferometric Measurements of Binary Stars 1.6Mb 12326 records

VI/46/

B-I-N-T McAlister H.A., HartkopF W.I., CHARA (Center for High Angular Resolution Astronomy, Georgia State University), Contr. 2 (1988 October Version) (1988)

Id.: WDS, HR, DM, ADS, name, HD epoch, ρ, θ, errors References and Notes By exchange from NASA-Goddard supersedes VI/24/

EVOLUTIONARY SPECTRA OF GALAXIES 1.7Mb 8 files

VI/47/

B-I-M-N-T-f ROCCA-VOLMERANGE B., GUIDERDONI B.: 1988, Astron. Astrophys. Suppl. Ser. 75, 93 GUIDERDONI B., ROCCA-VOLMERANGE B.: 1987 , Astron. Astrophys. 186, 1 λ, spectrum 8 files : burst, UV-cold and UV-hot E/SO, Sa, Sb, Sc, Sd, Im

Catalogue  References  Contents  Remarks

FeII Reference Catalogue 69Kb 645 references

VI/48/

B-I-M-N-T-f Viotti R., Baratta G.B., Astron. Astrophys. Suppl. Ser. 75, 497 (1988) Viotti R., Vittone A. Friedjung M. (eds.) , ”Physics of formation of FeII lines outside LTE” (IAU Coll. 94), D. Reidel Publ. Co., Dordrecht (1988)

Table 1: FeII reference catalogue Table 2: subject index Table 3: astrophysical categories see also Bull. Inform. CDS 36

Constellation Boundary Data 528Kb 1566 records

VI/49/

T-f-v Davenhall A.C., Leggett S.K., (1989)

Constellation Boundaries for 1875 and 2000 α, δ (J2000) σ

See also VI/42

Mean Energy Distributions in the Stellar Spectra of different Spectral Types and Luminosities. V. Normal Stars 159Kb 98 spectra

VI/50/

T-f-v Sviderskiene Z., Bull.Vilnius Obs. 80 (1988) λ (Δλ = 50Å) Intensities per unit wavelength interval By exchange from Moscow Data Center

COLLISIONAL EXCITA-TION OF FORMALDEHYDE IN “HOT” INSTERSTELLAR MOLECULAR REGIONS 354Kb 21 × 1600 rates

VI/51/

N-T-f-v GREEN S.: 1991, Astrophys. J. Suppl. Ser. 76, 979 Levels of ortho- and para-H_{2}CO Rates for ortho- and para-H_{2}CO levels for 21 temperatures between 10 and 300^{∘}K by exchange from NASA-Goddard

A LIST OF FAINT PHOTO-METRIC SEQUENCES (1980-1990) BASED ON PHOTOELECTRIC AND CCD MEASUREMENTS 111Kb 775 sequences

VI/53/

N-T RITZMANN B.M.: 1992 , Bull. Inf. CDS 40, 41 Field α, δ Detectors, colours mag. range, sources

Oxygen-enhanced evolutionary calculations for globular cluster stars. III. Horizontal-Branch sequences 594Kb 13 files

VI/54/

B-N-T-f-v Dorman B. , Astrophys. J. Suppl. Ser. 81, 221 (1992)

Age Y (He central fraction) log L, log T_{e}, log g log T_{c}, log ρ_{c}, log R M_{v}, B − V Masses 10 evolutionary sequences, and 55 ZAHB models

Catalogue  References  Contents  Remarks

EVOLUTIONARY TRACKS AND ISOCHRONES 15Mb 299 isochrones

VI/55/

N-T-f VANDENBERG D.A.: 1985, Astrophys. J. Suppl. Ser. 58, 711 & BELL R.A.: 1985, Astrophys. J. Suppl. Ser. 58, 561 ; Astrophys. J. Suppl. Ser. 63, 335 & LASKARIDES P.G.: 1987, Astrophys. J. Suppl. Ser. 64, 103 & PROFFITT C.R.: 1991, Astrophys. J. Suppl. Ser. 77, 473 VANDENBERG D.A.: 1992, Astrophys. J. 391, 685 & BERGBUSH P.A.: 1992, Astrophys. J. Suppl. Ser. 81, 163 Composition: Y Z [O/Fe] M, Age, log L, log T_{e}, log g Color indices for some isochrones Program to compute Luminosity Function ——————————— – Mass range 0.15 to 3 M_{⊙} Z range 0.0001 to 0.10

TESTS OF SHOCK CHEM-ISTRY IN IC 443G 71Kb 2×420 rates

VI/56/

N-T-f-v TURNER B.E., CHAN K.- W., GREEN S., LUBOWICH D.A., Astrophys. J. 399, 114 (1992)

Collisional excitation rates for CS − H_{2} and SiO − H_{2} rotational levels J = 0 ⋯ 20 and temperatures 20 to 300K by exchange from NASA-Goddard

Catalogue of Radii of Be star Line Emitting Regions 36Kb 119 stars

VI/57/

N-T-f-v Jaschek C., Jaschek M. , Astron. Astrophys. Suppl. Ser. 97, 807 (1993)

Id.: HD, HR

λ, Separation of emission line peaks Obs. date, reference v. sin i Radius of line emitting region

BIBLIOGRAPHIC CATALOGUE OF SUSPECTED VARIABLE STARS UPDATE 4.0Mb 88508 records

VI/58/

B-N-T RÖSSIGER S., BRÄUER H.J.: 1993 , Bull. Inf. CDS 42, 31 Id.: NSV Bibliography Keywords by exchange from Potsdam See also VI/52/

IMPROVED COLLISIONAL EXCITATION RATES FOR INTERSTELLAR WATER 387Kb 2 × 1980 rates

VI/59/

N-T-f GREEN S., MALUENDES S., McLEAN A.D.: 1993 , Astrophys. J. Suppl. Ser. 85, 181 Initial & Final H_{2}O levels Rates for ortho- and para-H_{2}O in collisions with He atoms for 11 temperatures between 20 and 2000^{∘}K by exchange from NASA-Goddard

CATALOGUE D’ARCHIVES DES PLAQUES PHOTOGRAPHIQUES DE L’OBSERVATOIRE D’ALGER 302Kb 5165 plates

VI/60/

N-T-f BOUNATIRO L., HANNI M., CHEMESE M.: 1993, Bull. Inf. CDS 43, 11 Id.: Plate number α, δ of plate centre Observation date Plate class

Catalogue  References  Contents  Remarks

Photoelectric Absorption cross-sections with variable abundances 48Kb 4 FORTRAN files

VI/62/

T-f Balucińska-Church M., Mc-Cammon D., Astrophys. J. 400, 699 (1992)

4 FORTRAN-77 files for computation of photoelectric and effective absorption cross-sections for most important elements between H and Ni; and interstellar photoelectric absorption by exchange from NASA-Goddard

ORBITAL, PRECESSIONAL, and INSOLATION QUANTI-TIES FOR THE EARTH from -20 Myr to +10Myr 7.2Mb +FORTRAN programs

VI/63/

N-T-f LASKAR J., JOUTEL F., BOUDIN F.: 1993, Astron. Astrophys. 270, 522 ϵ(t), e(t), ϕ(t) (nominal La90 solution of Earth orbit) Set of FORTRAN programs to compute the parameters of Earth orbit and climatic variables

Recombination line intensities for hydrogenic ions. IV. Total recombination coefficients and machine-readable tables for Z=1 to 8 154Mb 178 datafiles

VI/64/

N-T-f-v Storey P.J., Hummer D.G., Mon. Not. R. Astron. Soc. 272, 41 (1995)

162 primary output files with emissivities, effective recombination coefficients, line opacitiy factors for ions Z ≤ 8, log N_{e} = 2(1)14 16 “compressed intensity files” with an interactive data server Elements H, He, Li, Be, B, C, N, O

Ultraviolet radiation from evolved stellar populations. I. Models 2.0Mb 136 tracks

VI/65/

N-T-f-v Dorman B., Rood R.T., O’Connell R.W., Astrophys. J. 419, 596 (1993)

Grids for HB evolution: Y_{HB}, [F e/H], [O/F e] Mass, Age log L, Y_{c}, log T_{e}, log g log T_{c}, log ρ_{c}, log R

NUMERICAL EX-PRESSIONS FOR PRECES-SION FOR-MULAE AND MEAN ELE-MENTS FOR THE MOON AND THE PLANETS. 62Kb 3 Fortran programs

VI/66/

N-T-f SIMON J.L., BRETAGNON P., CHAPRONT J., CHAPRONT-TOUZE M., FRANCOU G., LASKAR J.: 1994, Astron. Astrophys. 282, 663 • Table of precession quantities BDL94

• Fortran77 codes to compute precession quantities and approximate epehmerides of the 8 major planets

Bibliographic Catalogue of Variable Stars (BCVS), Part I (1992 Edition) 17Mb 346570 entries

VI/67/

N-T-f Huth H., Wenzel W., Bull. Inf. CDS 20, 105 (1981)

Id.: Variable name Bibliography until 1987 Supersedes VI/35 & VI/38

Catalogue  References  Contents  Remarks

Bibliographic Catalogue of Variable Stars (BCVS), Part II Update 1994 8.5Mb 184064 entries

VI/68/

N-T-f-v RÖssiger S., Bräuer H.-J., Bull. Inf. CDS (unpublished)

Id.: Variable name

Bibliography

1982–1994 (and a few older references) Keywords Supersedes VI/52/ See also VI/67/

Atomic Spectral Line List 100Mb 629771 lines

VI/69/

T-f-v Hirata R., Horaguchi T., Department of Astronomy, Faculty of Science, Kyoto University and National Science Museum, 3-23-1 Hyakunin-cho, Shinjuku-ku, Tokyo Ion number, λ hybridation, multiplet log g_{f}, Intensity E, g References Remarks by exchange from NAO-Japan

A revised version of the ILLSS Catalogue 2.3Mb 25798 lines

VI/71A/

T-f-v COLUZZI R: 1993 , Bull. Inf. CDS 43, 7 λ, ion Multiplet numbers Intensity or χ or stellar types indications (revised version of Moore’s table)

Announced in Bull. 43 as III/171/ See also VI/84/

ATOMIC TRAN-SITION PROBABILITIES, Sc to Ni 3.6Mb 20018 transitions

VI/72/

T-f-v Martin G.A., Fuhr J.R., Wiese W.L. , J. Phys. Chem. Ref. Data, 17, Suppl. 3 (1988) Fuhr J.R., Martin G.A., Wiese W.L., J. Phys. Chem. Ref. Data, 17, Suppl. 4 (1988)

Element & Ionization Atomic Configurations Multiplet number λ, E_{low}, E_{up} g_{i}, g_{k}, A_{ik}, log gf References by exchange from NASA-Goddard See also VI/73/ and VI/74

Atomic Spectral Lines Data OII, Mg, Al, S, Sc 3.0Mb 14848 lines

VI/73/

T-f-v Martin W.C., Kaufman V.; Zalubas R., Musgrove A., Sugar J. , J. Phys. Chem. Ref. Data, 20, 83 (1991); 20, 775 (1991); 19, 821 (1990); 17, 1679 (1988)

Element & Ionization Atomic Configurations Multiplet number λ, E_{low}, E_{up} Relative intensity References by exchange from NASA-Goddard See also VI/72/ and VI/74

ATOMIC ENERGY LEVEL DATA 12Mb 47334 configurations

VI/74/

T-f-v Martin W.C., Kaufman V., Zalubas R., Musgrove A., Sugar J., Corliss C., J. Phys. Chem. Ref. Data, 10, 195 (1981); 9, 1 (1980); 8, 817 (1979); 12, 323 (1983); 14, 751 (1985); 19, 821 (1990); 14, Suppl.2 (1985); 19, 527 (1990); 20, 859 (1991); 17, 155 (1988) Martin W.C., Zalubas R., Hagan L., Natl. Stand. Ref. Data Ser., Natl. Bur. Stand.(U.S.) 60 (1978)

Element & Ionization Atomic Configurations Landé g factor by exchange from NASA-Goddard See also VI/72/ and VI/73

Simulation Atlas of Tidal Features in Galaxies 31Mb 84 simulations

VI/76/

f-v Howard S., Keel W.C., Byrd G.G., Burkey J., Astrophys. J. 417, 502 (1993)

168 FITS images, one gas + one star image for each encounter; each FITS file contains 1000 time steps by exchange from NASA-Goddard

Catalogue  References  Contents  Remarks

THEORETICAL STEL-LAR FLUX SPECTRA for Fto K-TYPE STARS 1.8Mb 242 models

VI/78/

T-f BUSER R., KURUCZ R.L., Astron. Astrophys. 264, 557 H_{ν} and H_{ν}^{cont} for 242 models with ranges: 3750K ≤ T_{eff} ≤ 6000K 0.75 ≤ log g ≤ 5.25 −3.00 ≤ [F e/H] ≤ +0.50

LUNAR SOLUTION ELP 2000-82B 3.7Mb 36 data files

VI/79/

T-f CHAPRONT-TOUZE M., CHAPRONT J.: 1983, 1988, Astron. Astrophys. 124, 50; 190, 342 Postscript description & Explanations 36 data files and a FOR-TRAN program to compute lunar ephemeris From Bureau des Longitudes

OPACITIES FROM THE Opacity Project 31Mb 213 mixtures

VI/80/

T-f SEATON M.J., YAN Y., MIHALAS D., PRADHAN A.K., Mon. Not. R. Astron. Soc. 266, 805 (1994) Mon. Not. R. Astron. Soc. 279, 95 (1996)

\underline{Codes} for reading and interpolating opacities to (T_{eff} , log g), to (X, Z) \underline{Data files}: 213 files with X− (hydrogen) and Z (“metals”) composition in solar metal-mix S92 96 files for interpolation in X and Z

See also VI/89

PLANETARY SOLUTIONS VSOP87 35Mb 48 data files

VI/81/

N-T-f BRETAGNON P., FRAN-COU G, Astron. Astrophys. 202, 309 (1988) BRETAGNON P., Astron. Astrophys. 114, 278 (1982)

48 data files used to compute elements and variables for planets (excluding Pluto) FORTRAN example program Documentation file From Bureau des Longitudes

Stark broadening of Hydrogen lines in the conditions of stellar enveloppes 2.0Mb 21822 records

VI/82/

T-f Stehle C., Astron. Astrophys. Suppl. Ser. 104, 509 (1994) Vienna International Workshop on Model Atmospheres and Spectrum syntheses for mid B to mid G stars, at or close to the Main Sequence (July 6 to July 11 1995) (to be pulished in Conferences Series of ASP)

10^{4} < T < 8 × 10^{4} 3.2 × 10^{14} < N_{e}(cm^{−3}) < 3.2 × 10^{16} intensities with Doppler or Stark only with H^{+}, He^{+}, Ar^{+} perturbers Concerns Lyman, Balmer, and Paschen lines Supersedes J/A+AS/104/509

UVILLSS Catalogue 1.6Mb 18431 lines

VI/84/

T-f-v Coluzzi R., Bull. Inf. CDS 48, 15 (1996) λ, ion Multiplet numbers Intensity ; χ

See also VI/71/

Ephemerides of planets between 1900 and 2100 2.6Mb FORTRAN program

VI/87/

T-f Chapront J., Francou G., Astron. Astrophys. 109, 181 (1995)

FORTRAN program and data for the computation of the Planetary Ephemerides, including Pluto See also VI/66 and VI/81 From Bureau des Longitudes

Catalogue  References  Contents  Remarks

Ephemerides of Pluto between 1900 and 2100 51Kb FORTRAN program

VI/88/

T-f Chapront J., Francou G., Astron. Astrophys. 109, 181 (1995)

FORTRAN program to compute the ephemerides of Pluto From Bureau des Longitudes

Radiative forces for stellar envelopes 163Mb Elements 06-28

VI/89A/

N-T-f Seaton M.J., Mon. Not. R. Astron. Soc. 279, 95, (1997) Mon. Not. R. Astron. Soc. 289, 700, (1997)

\underline{Codes} for computation of radiative accelerations, and Rosseland mean opacities \underline{Data files}: data for calculation of radiative accelerations for elements Z =6, 7, 8, 10, 11, 12, 13, 14, 16, 18, 20, 24, 25, 26 and 28

See also VI/80

Wide-Field Plate Database 38Mb 326635 plates

VI/90/

T-f-v Tsvetkov M.K., Stavrev K.Y., Tsvetkova K.P., Mutafov A.S., Semkov E.H., Institute of Astronomy, Bulgarian Academy of Sciences Id.: WFPDB α, δ (2000) Instrument aperture Observation date & time Observation method Object designation Observer Quality & availability Notes

Evolution Sequences of Stellar Models 7.1Mb 222 models

VI/96/

f-v Bressan A., Fagotto F., Bertelli G., Chiosi C., Astron. Astrophys. Suppl. Ser. 100, 647; 104, 365; 105, 29; 105, 39; 117, 113 (1993–1996)

Isochrones and Evolutionary Sequences for masses between 0.6 and 120M_{⊙} L, T_{eff}, g, T_{c}, X, Y, ^{12}C, 16O abundances as a function of age for 7 different metallicities

Solar and Lunar Eclipses 1996-2020 21Kb 114 eclipses

VI/97/

T-f-v Espenak F., Bull. Inf. CDS (in press)

Eclipse date & time (lon, lat), Path width Central duration

Extensive tabulations of Stark broadened hydrogen line profiles 791Kb 7922 profiles

VI/98A/

T-f-v , Astron. Astrophys. Suppl. Ser. 140, 93 (1999)

Id.: ; α, δ (J2000) ; σ Supersedes part of VI/82

Geneva Stellar Evolution Tracks and Isochrones 71Mb 4520 files

VI/102/

T-f-v Lejeune T., Schaerer D. , Astron. Astrophys. 366, 538 (2001)

Evolution tracks and Isochrones Z=0.0004 0.001 0.004 0.008 0.02 0.04 0.1 M ∈ [0.4 ⋯ 150]M_{⊙} Synthetic colours in UBV (RI)_{C}JKLMNH Geneva CMT_{1}T_{2} HST photometric systems

Catalogue  References  Contents  Remarks

HUT Catalog 117Kb 643 records

VI/103/

T-f-v Hopkins Ultraviolet Telescope Team , Johns Hopkins University (1990, 1995)

Id.: Target Observed α, δ (J2000) Date, exposure time Dataset identifiers

Via MAST

The Ultraviolet Imaging Telescope (UIT) 185Kb 1481 records

VI/104/

f-v Stecher T.P., et al., Astrophys. J. 395, L1 (1992); Publ. Astron. Soc. Pacific 106, 1151 (1994); Publ. Astron. Soc. Pacific 109, 584 (1997)

Id.: Target name α, δ (J2000) Exposure date, time Image number

The Wisconsin Ultraviolet Photo-Polarimeter Experiment (WUPPE) 61Kb 238 records

VI/105/

f-v Nordsieck K.H., Code A.D., Anderson C.M., Meade M.R., Babler B., Michalski D.E., Pfeifer R.H., Jones T.E. , Proc. SPIE Vol. 2010, p. 2-11 (1994)

Id.: Target name α, δ (J2000) Exposure Date and Time Polarisation qθ

Tübingen Ultraviolet Echelle Spectrometer 55Kb 229 records

VI/107/

f-v Grewing, M. (PI); Barnstedt J., et al., Astron. Astrophys. Suppl. Ser. 134, 561 (1999)

Id.: Target name α, δ (J2000) Exposure date, time Dataset names

Population Synthesis Models at very low metallicities 132Mb 255 models

VI/109/

f-v Schaerer D. , Astron. Astrophys. 397, 527 (2003)

Tracks: Z from Z = 10^{−7} IMF type Star formation history Colors of integrated population UV luminosities and colors

Final IUE Merged Observation Log 19Mb 110033 observations

VI/110/

f-v NASA / ESA , IUE Newsletters (1978-2000)

Id.: observer’s, homogenized α, δ (J2000) IUE Class Observation date+time Instrument+Image id. Comments from IUE-Vlispa Supersedes VI/99

ISO Observation Log 9.2Mb 36738 observations

VI/111/

f-v ISO Data Centre , ISO Data Centre (2004)

Id.: Target Name α, δ (J2000) Date/time Observation Mode, aperture Validation flag Links to ISO Images Supersedes VI/101

ISO Helpdesk:

helpdesk@iso.vilspa.esa.es

Catalogue  References  Contents  Remarks

Optical Imaging of 57 spiral galaxies 4040Mb 57 galaxies

VI/112/

f-v Knapen J.H., Stedman S., Bramich D.M., Folkes S.L., Bradley T.R. , Mon. Not. R. Astron. Soc. 344, 527 (2003); Astron. Astrophys. 426, 1135 Id.: NGC D_{25}, i, V_{r}, M_{B}, Distance Morph.Type, ArmClass Observational conditions Images in B, R, I, K, Hα

Evolutionary Models of Population II Stars 123Mb 285 tracks

VI/113/

f-v Cariulo P., Castellani V., Degl’Innocenti S. , Astron. Astrophys. 421, 1121 (2004)

Grids of Models Z = {2, 4, 6, 10, 40, 80} × 10^{−4} M ∈ [0.7 ⋯ 11]M_{⊙} \underline{Files} with: tracks, isochrones, HB-stage, structure parameters

Plate Centers of the second epoch Palomar Oschin Schmidt Sky Survey (POSS2) 337Kb records

VI/114/

f-v Read M. , Royal Obs. Edinburgh Id.: Plate, Field α, δ (J2000) of plate centers Observation dates Image qualities

The Digitized First Byurakan Survey (DFBS) plate database

184Kb 1729 plates

VI/116/

f-v Mickaelian A.M., Hagen H.-J., Sargsyan L.A., Mikayelyan G.M., Byurakan Observatory (2005)

Id.: FBS α, δ (B1950, J2000) Observation date Prism, emulsion Observer Quality

Observation  Log  from

ASCA 345Kb records

VI/117/

f-v

Id.: Observation ID, Target name α, δ (J2000) Observation Date Exposure times PI name

Stellar Models until He burning 936Kb 290 models

VI/118/

f-v Claret A., Gimenez A. , Astron. Astrophys. Suppl. Ser. 109, 441 (1995); Astron. Astrophys. Suppl. Ser. 114, 549 (1995); Astron. Astrophys. Suppl. Ser. 125, 439 (1997); Astron. Astrophys. Suppl. Ser. 133, 123 (1998)

Evolution models giving L g T_{eff} and apsidal motion as a function of time Z=0.03 0.02 0.01 0.004 Masses [1...40]M_{⊙}

A high accuracy computed water line list - BT2 12360Mb 505,806,202points

VI/119/

f-v Barber R.J., Tennyson J., Harris G.J., Tolchenov R.N. , Mon. Not. R. Astron. Soc. 368, 1087 (2006)

Id.: level# Transitions: A FORTRAN program σ

Catalogue  References  Contents  Remarks

Library of synthetic stellar spectra (0.3-1.8μm) 10076Mb 2705 files

VI/120/

f-v Coelho P., Barbuy B., Melendez J., Schiavon R., Castilho B. , Astron. Astrophys. 443, 735 (2005)

2705 FITS files with spectra in range 3000 − 18000Å, Δλ = 0.02Å T_{eff} = 3500–7000K, ΔT =250K log g = 0–5.0, Δ log g=0.5 [Fe/H] = −2.5, −2.0, −1.5, −1.0, −0.5, 0.0, 0.2, 0.5 [α/Fe] = 0.0, 0.4

Energy level list and linelist of HCN/HNC 3538Mb 34433190 lines

VI/121/

f-v Harris G.J., Tennyson J., Kaminskiy B., Pavlenko Y.V., Jones H.R.A. , Mon. Not. R. Astron. Soc. 367, 400 (2005)

Energy levels Transitions: quantum numbers, ν

See also VI/119

The Marginal Zone of the Moon 116Mb records

VI/122/

f-v Watts C.B. , Astron. Papers for the preparation of the Nautical Almanach XVII (1963)

\underline{Moon limb}: height, accuracy \underline{Moon photographs}: date, libration, measured arcs

Theoretical near-IR spectra of red supergiants and giants 127Mb 403 spectra

VI/125/

f-v Lançon A., Hauschildt P.H., Ladjal D., Mouhcine M. , Astron. Astrophys. 468, 205 (2007)

Id.: Model M, T_{eff}, log g Spectra 0.51– 2.49μm, Δλ=0.25Å(R ∼ 40000) for mass M = {1M, 15}M_{⊙}, composition ⊙ + RSG

VII. Non-stellar and Extended Objects

Catalogue  References  Contents  Remarks

The Revised New General Catalogue of nonstellar astronomical objects 1.2Mb 7840 records

VII/1B/

L-M-N-T-f-v Sulentic J.W., Tifft W.G., Univ. of Arizona Press (1974)

Id.: RNGC α, δ (1975) ; ℓ, b x, y position on POSS, m Old and New Descriptions Cross-identifications

A Master list of Radio Sources

7.1Mb 79494 objects

VII/2A/

N-T-f-v Dixon R.S., Astrophys. J. Suppl. Ser. 20, 1 (1970, updated 1978)

Identification α, δ (1950) ν, F (ν) Reference

Catalogue of Abell and Zwicky Clusters of galaxies 1.9Mb 21179 records

VII/4A/

T-f-v Abell G.O., Astrophys. J. Suppl. Ser. 3, 211 (1958) Corwin H.G., Astron. J. 79, 1356 (1974) Zwicky F., Herzog E., Wild P., Karpowicz M., Kowal C.T., Catalogue of galaxies and of clusters of galaxies, California Institute of Technology, Pasadena, 6 vols. (1961-68)

Abell clusters Zwicky clusters Cross-identifications for Zwicky clusters Tape prepared by the Department of Astronomy of the Bulgarian Academy of Science, Sofia, Bulgaria

EXCERPT FROM THE CAT-ALOGUE OF STARS CLUS-TERS AND ASSOCIATIONS by G. ALTER ET AL. 83Kb 1039 records

VII/5A/

I-L-M-N-T-f-v KLUKE J., MOFFAT A.F.J., SCHMIDT-KALER TH., VOGT N.: 1975, Astron. Institut, Ruhr-Univ. Bochum Id.: Cluster Name α, δ(1950) ℓ, b Cluster type Distance, Diameter, etc for all open clusters known in 1968 from the Catalogue of Stars Clusters & Associations by G. Alter, J. Ruprecht, V. Vanysek (Budapest 1970)

CD-ROM Version See also VII/31/, VII/44/, VII/101

CATALOGUE OF POLAR-IZATION MEASUREMENTS AND RELATED DATA OF EXTRA-GALACTIC RADIO SOURCES 188Kb 510 sources

VII/6/

N-T-f-v EICHENDORF W., REIN-HARDT M.: 1979 , Astrophys. Sp. Sci. 61, 153 Acta Cosmologica 9, 1 Bull. Inform. CDS 15, 122 Id.: 3C αδ (1950.0) Classification Polarization angle, density Diameter, Rotation Redshift z Spectral indices Polarization indices

Lynds’ catalogue of dark nebulae 114Kb 1791 nebulae

VII/7A/

M-N-T-f-v Lynds B.T., Astrophys. J. Suppl. Ser. 7, 1 (1962) (Updated version)

Id.: Lynds, Barnard α, δ (1950) ; ℓ, b Area (square degrees) Opacity By exchange from NASA-Goddard See also VII/191/

Catalogue  References  Contents  Remarks

Lynds’ Catalogue of Bright Nebulae (LBN) 75Kb 1125 nebulae

VII/9/

T-f-v Lynds B., Astrophys. J. Suppl. Ser. 12, 163 (1965)

Id.: NGC, IC, S, C, DG α, δ (1950) ; ℓ^{II} , ℓ^{II} Dimensions & area Color Brightness By exchange from NASA-Goddard.

Kitt Peak Quasar List 13Kb 261 quasars

VII/11/

B-M-N-T-f-v Deveny J.B., Osborn W.H., Janes K., Publ. Astron. Soc. Pacific 83, 611 (1971)

{Id.: Name \atop α, δ (1950)}

Selected Data only

Arp’s Globular Cluster Catalog 24Kb 119 records

VII/13/

B-M-N-T-f-v Arp H.C., in: Stars and Stellar Systems, Blaauw A., Schmidt M. (eds.), 5, 401 (1965)

Od.: Cluster Name

α, δ (1950)

By exchange from NASA-Goddard See also VII/151/

Galactic Supernova Remnants Catalog 14Kb 120 SNRs

VII/14/

B-M-N-T Clark D.H., Caswell J.L., Mon. Not. R. Astron. Soc. 174, 267 (1979)

Id.: SNR

F (408MHz), F (5GHz) Spectral index Angular Diameter Surface brightness References By exchange from NASA-Goddard

Galactic Supernova Remnants Catalog 13Kb 116 SNRs

VII/15/

B-M-N-T Ilovaisky S.A., Lequeux J., Astron. Astrophys. 18, 169 (1972)

Id.: SNR

ℓ, b

F (1GHz)

Angular diameter Surface brightness Diameter, distance, R, z References By exchange from NASA-Goddard

FOURTH UHURU CATALOGUE 20Kb 339 records

VII/18/

B-M-N-T FORMAN W., JONES C., COMINSKY L., JULIEN P., MURRAY S., PETERS G., TANANBAUM H., GI-ACONNI R.: 1978, Astrophys. J. Suppl. Ser. 38, 357 Id.: 4U α, δ (1950) X-ray intensity Comments By exchange from NASA-Goddard

Second Ariel Catalogue . 107 records

VII/19/

B-M-N-T Cooke B.A., Ricketts M.J., MacCacaro T., Pye J.P., Elvis M., Watson M.G., Griffiths R.E., Pounds K.A., McHardy I., Maccagni D., Seward F.D., Page C.G., Turner M.J.L., Mon. Not. R. Astron. Soc. 182, 489 (1978)

Id.: 2A α, δ (1950) X-Ray intensity Comments By exchange from NASA-Goddard

Catalogue of HII regions 31Kb 313 records

VII/20/

B-M-N-T-f-v Sharpless S., Astrophys. J. Suppl. Ser. 4, 257 (1959)

Id.: Sharpless

α, δ (1860, 1900, 1950)

ℓ^{II} , b^{II}

Angular diameter

Form, Type, brightness Number of associated stars By exchange from NASA-Goddard

Catalogue  References  Contents  Remarks

Catalogue of Reflection Nebulae 19Kb 158 nebulae

VII/21/

T-f-v van den Bergh S., Astron. J. 71, 990 (1966)

Id.: HD, BD ℓ , b Sp, m_{V} Type, surface brightness Colors, Absorption Max radius By exchange from NASA-Goddard

A Catalogue of absorption lines in QSO spectra 492Kb 108 records

VII/23/

B-M-N-T Ellis R., Mon. Not. R. Astron. Soc. 185, 613 (1978)

Id.: Name α, δ (1950) UBV z(emission) Reference Line list

Uppsala General Catalogue of Galaxies (UGC) 1.0Mb 12940 galaxies

VII/26D/

T-f-v

[Identifiers in Simbad]

Nilson P., Uppsala Astron. Obs. Ann. 6 (1973)

Id.: UGC, MCG, PSS α, δ (1950) Dimensions, i Classification m, V_{r} Updated ADC CD-ROM version

Catalog of Southern Groups and Clusters of Galaxies 51Kb 967 clusters

VII/28/

N-T-f-v Duus A., Newell B., Astrophys. J. Suppl. Ser. 35, 209 (1977)

Id.: Number, ESO, SRC Distance of center Concentration mass N(galaxies) Distance class References Tape version produced by G. Share By exchange from NASA-Goddard

CATALOGUE OF STAR CLUSTERS AND AS-SOCIATIONS. Part I: ASSO-CIATIONS 325Kb 89 associations

VII/31B/

N-T-f-v RUPRECHT J., BALAZS B., WHITE R.E.: 1982, Bull. Inform. CDS 22, 132 (supplement up to 1973 of the Catalogue of Star Clusters and Associations by Alter G., Ruprecht J., Vanysek V., Budapest 1970)

Bibliographical entries and basic information on each individual object.

See also VII/5/, VII/44, VII/101/ By exchange from Moscow Data Center

CATALOGUE OF HIGH REDSHIFT 402Kb 11864 records

VII/32/

B-M-N-T FLICHE H.H., SOURIAU J.M., TRIAY R.: 1982 , Astron. Astrophys. 108, 256 \underline{First file}: α, δ, FC ref. z, UBV Polarisation, Absorption References \underline{Second file}: Object type, names, references comments chemical composition

The ESO/Uppsala Survey of the ESO (B) Atlas 2.4Mb 18438 objects

VII/34C/

T-f-v

[Identifiers in Simbad]

Lauberts A., European Southern Obs. (1982)

Id.: ESO, other α, δ (1950) ; ℓ, b; (x, y) Object class, Dimensions Classification, UBV, V_{r} Description Notes and References Plates history See VII/115/

Catalogue  References  Contents  Remarks

CATALOG OF GALAXY REDSHIFTS 330Kb 3981 galaxies

VII/36/

N-T-f-v ROOD H.J.: 1980, (unpublished)

Id.: Number, UGC, NGC, IC α, δ ; ℓ, b m_{B} , z, V_{r}, σ Source By exchange from NASA-Goddard

DETAILED BIBLIOG-RAPHY OF THE SURFACE PHOTOMETRY OF GALAX-IES 231Kb 702 galaxies

VII/39A/

N-T-f-v DAVOUST E., PENCE J.D., Astron. Astrophys. Suppl. Ser. 49, 631 (1982)

Id.: NGC, IC

UBV RI emission lines Luminosity profile References

A Catalogue of Magnitude, Colours, Ellipticities and Position Angles for the field of Coma Cluster 699Kb 6724 records

VII/42A/

N-T-f-v Godwin J.G., Metcalfe N., Peach J.V., Mon. Not. R. Astron. Soc. 202, 113 (1983)

Id: GMP, GP, RB67 X, Y ; b_{26.5}, r_{24.75} r^{∗}(b_{26.5}), r^{∗}(r_{24.75}) ellipticity ε and ϑ α_{b} , γ_{b} ; Comments

WISCONSIN SOFT X-RAY DIFFUSE BACKGROUND ALL-SKY SURVEY 11Mb 40 files

VII/43/

B-M-N-T McCAMMON D., BUR-ROWS D.N., SANDERS W.T., KRAUSHAAR W.L.: 1983, Astrophys. J. 269, 107 Utility Subroutines Maps and count-rates in soft X-ray bands

CATALOGUE OF STAR CLUSTERS AND ASSOCIATIONS. II. GLOBU-LAR CLUSTERS 954Kb 137 clusters

VII/44B/

B-L-M-N-T-f-v RUPRECHT J., BALAZS B., WHITE R.E.: 1983, Soviet Astronomy 27, 358 Bibliographical entries and basic information on globular clusters only.

See also VII/5/, VII/31/, VII/101/ By exchange from Moscow Data Center

CATALOGUE OF MIDDLE UV PHOTOM-ETRY OF VIRGO CLUSTER GALAXIES 18Kb 201 records

VII/48/

B-I-M-N-T-f CORNETT R.H., SMITH A.M.: 1981, 1982 , Astrophys. J. 261, 1 A.D.C. Bull. 1, 78 Id.: NGC, Anon, IC Revised Hubble type axial ratio by exchange from NASA-Goddard

CO RADIAL VELOCITIES TOWARD GALAC-TIC HII REGIONS: CO OB-SERVATIONS OF SOUTH-ERN HEMISPHERE HII RE-GIONS 94Kb 398 sources

VII/50/

N-T-f-v BLITZ L., FICH M., STARK A.A.: 1982, Astrophys. J. Suppl. Ser. 49, 183 GILLESPIE A.R., HUG-GINS P.J., SOLLNER T.C.L.G., PHILLIPS T.G., GARD-NER F.F., KNOWLES S.H., Astron. Astrophys. 60, 221 Id.: CO α, δ ; ℓ, b V_{r}(CO), σ T_{A}, FWHM, diameter, comments Velocity of the H109α, OH and H_{2}CO lines References & Notes By exchange from NASA-Goddard

Catalogue  References  Contents  Remarks

Revised Shapley-Ames Catalog of Bright Galaxies 100Kb 1416 galaxies

VII/51/

B-M-N-T-f-v

Sandage  A.,  Tam-

mann G.A., Carnegie Institution of Washington Publication 635 (1981)

Id.: NGC, IC α, δ (1950) m, Type, luminosity class V_{r}, σ Internal absorption Cluster membership By exchange from NASA-Goddard

A Catalogue of Extragalactic Radio-Source Identifications 1.2Mb 14585 sources

VII/54/

B-L-M-N-T-f-v Véron-Cetty M.P., Véron P., Astron. Astrophys. Suppl. Ser. 53, 219 (1983)

Id.: Name, 4C Optical identification α, δ (1950) m_{v}, z Remarks References & Notes By exchange from NASA-Goddard

A TABLE OF REDSHIFTS FOR ABELL CLUSTERS 19Kb 431 clusters

VII/56/

N-T-f-v SARAZIN C.L., ROOD H.J., STRUBLE M.F.: 1982, Astron. Astrophys. 108, L7 Id.: Abell zReferences Notes By exchange from NASA-Goddard

NEAR-INFRARED SOURCE COUNTS IN THE GALAC-TIC PLANE 115Kb 1989 sources

VII/60/

B-N-T-f KAWARA K., KOZASA T., SATO S.: 1982 , Publ. Astron. Soc. Jap. 34, 389 Mem. Fac. Science, Kyoto Univ., Ser. A, 36, 353 α, δ (1950); ℓ, b IHKL magnitudes by exchange from NASA-Goddard

Morphological Catalog of galaxies (MCG) 2.0Mb 29003 galaxies

VII/62A/

B-N-T-f-v

[Identifiers in Simbad]

Vorontsov-Velyaminov B.A., Arkhipova V., Sternberg Inst. Moscow, Vols. I-IV (1962-1974)

Id.: MCG, NGC, IC α, δ m_{pg} , Inclination Dimensions (inner part and entire galaxy) Intensities (inner and outer regions)

A  few  errors

corrected from CD-ROM version see also VII/100/

A LIST OF UV EXCESS GALAXIES 17Kb 412 records

VII/63/

B-M-N-T MITCHELL K.J.,BROTZMAN L.E., WARNOCK A., USHER P.D.: 1982, Astrophys. Sp. Sci. 88, 219 α, δ m, color code Morphology Diameters By exchange from NASA-Goddard

Catalog of CO Observations of Galaxies 105Kb 948 records

VII/64/

B-I-M-N-T-f-v Verter F., Astrophys. J. Suppl. Ser. 57, 261 (1985)

Id.: galaxy name Morphological type Diameter, m_{B} V_{r}, F (HI), T_{A}, F (CO) References, Telescopes By exchange from NASA-Goddard detect uprlmits maps Ref

DARK NEBU-LAE AND GLOBULES FOR ℓ = 240^{∘} to 360^{∘} 2.2Mb 2210 records

VII/68A/

M-N-T-f-v FEITZINGER J.V., STÜWE J.A.: 1984, Astron. Astrophys. Suppl. Ser. 58, 365 Id.: Field-identity α, δ ; ℓ, b ; x, y Cross references Area, classification Inclination Sky map (FITS)

Catalogue  References  Contents  Remarks

The Asiago Catalogue of QSOs

274Kb 2004 QSOs

VII/69/

T-f-v

Barbieri  C.,  Capac-

cioli M., Custiani S., Nardo G., Omizzolo A., Mem S.A. It. 54, 601 (1983)

Id.: Name α, δ UBV Photometry z References & Remarks

Supersedes VII/3/

GALAXIES ROTATION CURVES: A CATALOGUE 59Kb 931 records

VII/70A/

B-M-N-T-f-v BAIESI-PILLASTRINI G.C., PALUMBO G.G.C., VETTOLANI G.: 1983 , Astron. Astrophys. Suppl. Ser. 53, 373 Id.: NGC, other α, δ (1950) Morphological type, V_{r} Extent of the rotation curve

Reference

Redshifts for Zwicky’s Near Clusters 38Kb 503 clusters

VII/71/

B-N-T-f-v Baiesi-Pillastrini G.C., Palumbo G.G.C., Vettolani G. , Astron. Astrophys. Suppl. Ser. 56, 363 (1984)

Id.: UGC, Zwicky α, δ (1950) Contour ; V_{r} N(galaxies) Notes

A Catalog of Stellar Velocity Dispersions. I. Compilation and Standard Galaxies 135Kb 724 galaxies

VII/72/

B-M-N-T-f-v Whitmore B.C., McElroy D.B., Tonry J.L., Astrophys. J. Suppl. Ser. 59, 1 (1985)

Id.: NGC Type Dispersion, FWHM Slit (width and length)

No coordinates

IRAS Small Scale Structure Catalog 3.3Mb 16740 sources

VII/73/

N-T-f-v Helou G., Walker D.W. (Joint IRAS Science W.G.), Jet Propulsion Laboratory (1985)

Id.: IRAS α, δ (1950) F (12, 25, 60, 100μm) Associations by exchange from NASA-Goddard

Atlas of Peculiar Galaxies 17Kb 338 records

VII/74A/

T-f-v ARP H.C.: 1966, Astrophys. J. Suppl. Ser. 14, 1 (Table 2)

Id.: Arp, NGC, IC α, δ (1970) Corrected z Keypunched at the CDS See also VII/192/

A Catalog of Infrared Photometry and 21 cm data for Nearby Galaxies 6.0Kb 307 galaxies

VII/75/

M-N-T-f-v Aaronson M., Huchra J., MoulD J.R., Tully R.B., Fisher J.R. , van Woerden H., Mebold U., Siegman B., Berriman G., Persson S.E., Astrophys. J. Suppl. Ser. 50, 241 (Table 3) (1982)

Id.: NGC, other α, δ (1950) Morph.Type, i H_{−0.5}^{c}, log D_{1}, V, ΔV_{20}^{c} F (HI), d/d_{V irgo} Group assignment References & Notes Keypunched at the CDS

Catalogue  References  Contents  Remarks

Catalogue of isolated Pairs of galaxies in the Northern Hemisphere 92Kb 1206 galaxies

VII/77/

M-N-T-f-v Karachentsev I.D., Lebedev V.S., Shcherbanovskij A.L., Publ. Special Astrophys. Obs. of USSR 7 (1972) Bull. Inform. CDS 29, 87 (1985)

Id.: K72 Number α, δ (1950) m_{v}, Sp V_{r}, σ, source Dimensions Morphological type Angular separation By exchange from Moscow Data Center

HEAO A-1 X-RAY SOURCE CATALOG 262Kb 842 records

VII/80/

N-T-f-v WOOD K.S., MEEKINS J.F., YENTIS D.J., SMATHERS H.W., McNUTT D.P., BLEACH R.D., BYRAM E.T., CHUBB T.A., FRIEDMAN H., MEIDAV M.: 1984, Astrophys. J. Suppl. 56, 507– 649 Id. : 1H, other α, δ(1950) (95% confidence)

ℓ^{II}, b^{II} + Ecliptic Area, F (0.5–25keV) by exchange from NASA-Goddard

Catalogue of Isolated Galaxies 61Kb 1051 galaxies

VII/82A/

M-N-T-f-v Karachentseva V.E., Lebedev V.S., Shcherbanovskij A.L., Comm. of Special Astrophys. Obs. USSR 8 (1973) Bull. Inform. CDS 30, 125 (1986)

Id.: Number α, δ (1950) Isolation class m_{v}, Morphological type Dimensions V_{r}, σ, source By exchange from Moscow Data Center

CATALOGUE OF ISOLATED TRIPLETS OF GALAXIES 25Kb 84 triplets

VII/83/

N-T-f-v KARACHENTSEVA V.E., KARACHENT-SEV I.D., LEBEDEV V.S., Astrofiz. Issled. 26, 42 Id.: Number α, δ (1950) m_{v}, dimensions Separation V_{r}, σ By exchange from Moscow Data Center

CATALOGUE OF GROUPS OF GALAXIES 18Kb 372 records

VII/84A/

N-T-f-v GELLER M.J., HUCHRA J.P.: 1983 , Astrophys. J. Suppl. Ser. 52, 61 (tables 1 & 3)

Id.: Number, CfA, Turner & Gott, Huchra & Geller, de Vaucouleurs, other α, δ (1950) N(galaxies) Corrected V_{r}, dispers. Integrated magnitude Mean harmonic radius Mean pairwise sep.

CATALOGUE OF NEARBY GROUPS OF GALAXIES 13Kb 124 records

VII/86A/

N-T-f-v HUCHRA J.P., GELLER M.J.: 1982 , Astrophys. J. 257, 423 (tables 4 & 5, col 1-7)

Id.: Number, other α, δ (1950) N(galaxies) V_{r}, dispersion Total magnitude

Data on 1989 Abell’s rich clusters of galaxies 58Kb 1889 records

VII/87A/

N-T-f-v Leir A.A., van den Bergh S., Astrophys. J. Suppl. Ser. 34, 381 (1977) (Table 1, cols 1,6,7,9-11)

Id.: Abell Bautz-Morgan class Apparent radius Apparent dimensions m_{R}(brightest) by exchange from NASA-Goddard

Catalogue  References  Contents  Remarks

CEDAG CAT-ALOGUE OF CLUSTERS OF GALAXIES 1.1Mb 10496 records

VII/88/

N-T-f-v FERNANDEZ A., MATHEZ G., NOTTALE L.: 1984,, Journées de Strasbourg, 6ème réunion, Strasbourg Obs., 33 Id.: CEDAG, Abell, Zwicky, Kalinkov α, delta (Abell, Zwicky) Supercluster Angular diameter (Zwicky)

Magnitude (Abell) Distance groups (Abell, Zwicky) Richness group (Abell) Population (Zwicky)

Catalogue of Compact Groups of Compact Galaxies 25Kb 377 groups

VII/89B/

N-T-f-v Shakhbazian R.K., Astrophysics 9, 296 (1973) Shakhbazian R.K., Petrosian M.B. , Astrophysics 10, 6 (1974) Baier F.W., Petrosian M.B., Tiersch H., Shakhbazian R.K., Astrophysics 10, 202 (1974) Petrosian M.B., Astrophysics 10, 291 (1974) Baier F.W.,Tiersch H., Astrophysics 11, 146; 12, 1 ; 12, 263 ; 14, 157 ; 15, 24 (1975-79) Petrosian M.B. , Astrophysics 14, 356 (1978)

Id.: Number, other α, δ (1950) Apparent dimension N(galaxies) m(brightest) Relative compacity See also VII/196

MERGED CATALOGUE OF GALAXIES 7.4Mb 31917 galaxies

VII/90A/

N KOGOSHVILI N.G.: 1975, 1981, 1983, Bull. Abastumani Astr. Obs. 46, 133 IAU Coll. 64, Reidel D. Publ. Co., 273 Bull. Inform. CDS 25, 63 Id.: MCG, UGC, NGC, IC α, δ (1950) Angular diameters Surface brightness Inclination, Dimensions Morphological types V_{r}, Rot. velocity SNs and HII regions UBV F (178, 780MHz) F (1.4, 2.7, 5GHz) Other components References Contains all galaxies with −33^{∘} < δ < +90^{∘} m < 15.1 by exchange from Moscow Data Center supersedes VII/40/

IRAS ASTEROID AND COMET SURVEY 6.0Mb 22949 records

VII/91/

M-N-T-f-v

JOINT  IRAS  SCI-

ENCE W.G.: 1986, MAT-SON D.L. (ed.)

Id.: IRAS Date of observation, α, δ F (12, 25, 60, 100μm) S/N, Variability Orbital elements Distances V, B − V Albedo, diameter.. by exchange from NASA-Goddard

A PALOMAR OBSERVATORY SKY SUR-VEY ATLAS OF SELECTED MOLECULAR CLOUDS 88Mb 43 FITS images

VII/94/

N-T-f KLINGLESMITH D.A., HOLLIS J.M.: 1987,, Astrophys. J. Suppl. Ser. 64, 127 Source name α, δ (1950) Flux density By exchange from NASA-Goddard

Catalogue  References  Contents  Remarks

A CATALOG OF MORPHO-LOGICAL PROPERTIES OF ABELL CLUSTERS 619Kb 2712 clusters

VII/96/

N-T-f-v STRUBLE M.F., ROOD H.J. , Astrophys. J. Suppl. Ser. 63, 555 (1987)

Id.: Abell RS type Galaxy Contents & Separations Notes & Comments

A CATALOGUE OF 2810 NEARBY GALAXIES 499Kb 8430 records

VII/98A/

T-f-v KRAAN-KORTEWEG R.C.: 1986, Astron. Astrophys. Suppl. Ser. 66, 255 Id.: NGC, IC α, δ Morph. type, Lum. class Dimensions, V_{r} m_{B}, absorption References Remarks

Catalogue of Radial Velocities of Galaxies 1.2Mb 8250 galaxies

VII/99/

T-f-v Palumbo G.G.C., Tanzellanitti G., Vettolani G., Gordon and Breach (Ed) (1986)

Id.: UGC, other α, δ (1950) V_{r}, σ(V_{r}) Reference number Comments

Compilation of the Fifth Volume of Morphological Catalogue of Galaxies (MCG) 440Kb 3518 records

VII/100/

T-f-v VORONTSOV-VELYAMINOV B.A., ARKHIPOVA V.: 1974, Moscow Univ.

Id. : MCG, NGC, IC α, δ (1950) m_{pg}, dimensions, V_{r} Surface brightness Morphological type SNs, HII regions UBV See also VII/62/ By exchange from Moscow Data Center

Catalogue of Star Clusters and Associations. III. Open Clusters 2.1Mb 1112 clusters

VII/101A/

N-T-f-v Ruprecht J., Balazs B., White R.E., Soviet Astronomy 27, 358 (1983)

Bibliographical entries and basic information on open clusters only.

See also VII/5/, VII/31/, VII/44/ By exchange from Moscow Data Center

Spectrophotometric Atlas of Synthesis Standard Spectra of early-type Galaxies 846Kb 96 FITS files

VII/102/

M-N-T-f-v Pickles A.J.: 1985, Astrophys. J. Suppl. Ser. 59, 33 Synthesis standard spectra for 48 common spectral types, at resolution of 3 and 12 Å/pixel

Catalogue of Galactic Globular Clusters 22Kb 160 clusters

VII/103/

B-M-N-T-f-v Monella R., Coelum LIII, 287 (1985)

Id.: Number, name α, δ (1950) m_{v}, M_{v}, Sp, [F e/H] Diameters, Distances, V_{r} m(brightest star) Concentration class N(variables) keypunched at the C.D.S. See also VII/151

Catalogue  References  Contents  Remarks

Open Cluster Interstellar Matter Database 1.4Mb 128 clusters

VII/106/

f-v Leisawitz D.T. , NASA Ref. Publ. 1202 (1988)

Id.: OCl, Sh, other ℓ, b (J2000) Various data on clusters References

Asteroids II Data Base Version March 1988 1.2Mb 6213 asteroids

VII/108C/

B-I-N-T-f-v Lagerkvist C.- I., Harris A.W., Zappala V., Univ. Arizona Press, Tucson (1987)

Id.: Asteroid names Discovery, Elements, family Lightcurve parameters Pole determinations Taxonomic classifications UBV, albedos Diameters Discovery files updated Dec-1990 By exchange from NASA-Goddard

A CATALOG OF IRAS OB-SERVATIONS OF LARGE OPTICAL GALAXIES 72Mb 83 galaxies

VII/109/

N-T-f-v RICE W., LONSDALE C.J., SOIFER B.T., NEUGE-BAUER G., KOPAN E.L., LLOYD L.A., DE YONG T., HABING H.J.: 1988, Astrophys. J. Suppl. Ser. 68, 91 664 FITS images 4 images per galaxy in the four IRAS passbands at 12, 25, 60 and 100μm, and the corresponding 4 noise arrays By exchange from NASA-Goddard

A Catalogue of Rich Clusters of Galaxies 1.0Mb 5524 clusters

VII/110A/

N-T-f-v Abell G.O., Corwin H.G., Olowin R.P., Astrophys. J. Suppl. Ser. 70, 1 (1989)

Id.: Abell, field α, δ (1950, 2000) ; (x, y) ℓ, b ; z, R, D Abell and Bautz-Morgan types N(gal.), m_{3}, m_{10}, m_{10a} Observers

* Notes By exchange from NASA-Goddard #A145 in H.Andernach’s list

The Second Reference Catalogue of Bright Galaxies (RC2) 1.3Mb 4364 galaxies

VII/112/

T-f-v de Vaucouleurs G., de Vaucouleurs A., Corwin H.G., University of Texas Press, Austin (1976)

Id. RNGC α, δ (1950) ; ℓ, b ; L, B Morphol. Type, T, L Nuclear type R_{25}, D_{25}, A_{e}, D_{0} DDO types m_{H}, B_{T}, m′_{e}, m′_{25} Color indices Galactic extinction Radio S_{R}, S_{H}, RI Radio indexes α^{+}, α^{−} V_{r}, σ, N(meas.)

By exchange from NASA-Goddard 2nd ed. of VII/16/

See VII/155/

Catalogued Galaxies and Quasars observed in the IRAS Survey, Version 2 3.3Mb 12776 galaxies

VII/113/

B-I-M-N-T-f-v IPAC : 1989,

2 FITS tables by exchange from NASA-Goddard

Catalogue  References  Contents  Remarks

Massachusetts-Stony Brook Galactic Plane CO Survey. II. (ℓ, V) maps of the First Galactic Quadrant 27Mb 3 FITS files

VII/114/

B-M-N-T-f-v Clemens D.P., Sanders D.B., Scoville N.Z., Solomon P.M. , Astrophys. J. Suppl. Ser. 60, 297 (1986)

File 1: ℓ = 8^{∘} − 17.9^{∘}
File 2: ℓ = 18^{∘} − 55^{∘}
File 3: ℓ = 55.1^{∘} − 89.9^{∘}

by exchange from NASA-Goddard

The surface photometry catalogue of the ESO-Uppsala galaxies 49Mb 15467 galaxies

VII/115/

B-I-N-T-f-v Lauberts A., Valentijn E.A., ESO (ISBN 3-923524-31-5) (1989)

Id.: ESO, NGC, IC α, δ (1950) ; ℓ, b V_{r}, TType Integrated and Surfaces magnitudes Dimensions Profile shapes Environmental parameters

See VII/34/

Southern Galaxy Catalogue 1.4Mb 5472 galaxies

VII/116/

N-T-f-v Corwin H.G., de Vaucouleurs A., de Vaucouleurs G., The University of Texas Monographs in Astronomy No.4, The Department of Astronomy, University of Texas at Austin (1985)

Id.: RC2, Field-EU α, δ (1950, 2000), ℓ, b Morph.Type, num. type Weighted mean lum. class,

log D, log R + weights

Extensive notes by exchange from NASA-Goddard supplied by H.Andernach

NGC 2000.0 1.3Mb 13226 objects

VII/118/

T-f-v DREYER J.L.E., “The Complete New General Catalogue and Index Catalogue of Nebulae and Star Clusters by J.L.E. Dreyer ”, edited by Sinnott R.W., 1988, Sky Publication Corporation and Cambridge University Press Id.: NGC or IC α, δ (2000) m, dimensions Summary description copyrighted: data distributed for research purposes only by exchange from NASA-Goddard

COS-B FINAL DATABASE 24Mb 20 files

VII/120/

B-I-M-N-T-v MAYER-HASSELWANDER H.A.: 1985, Explanatory Supplement to the COS-B final database, issue 1, October 1985, The caravane collaboration for the COS-B satellite Data and related information experiment characteristic functions Programs for access to database Skymaps of the galactic disc

CO Observations of Galaxies 1985-1989 52Kb 142 observations

VII/139/

T-f-v Verter F., Publ. Astron. Soc. Pacific 102, 1281 (1990)

Name of 1st author Year of publication Telescopes Transistions measured Source selection by exchange from NASA-Goddard

A Spectrophotometric Atlas of Galaxies 793Kb 55 galaxies

VII/141/

N-T-f-v Kennicutt R.C.Jr, Astrophys. J. Suppl. Ser. 79, 255 (1992)

Id.: NGC, Mkr F (λ)/F (5500Å) 1726 points from 3600 to 7050Å, step 2Å by exchange from NASA-Goddard

Catalogue  References  Contents  Remarks

Southern Redshifts Catalogue 1.2Mb 12844 galaxies

VII/142A/

B-N-T-f-v Fairall F.P., Jones A , Publ. Department Astron., Univ. Cape Town 11 (1991)

Id.: ESO, NGC, other α, δ (1950) ℓ, b V = cz (heliocentric) Remark , references Galaxy type

Supersedes VII/107/

Quasar Candidates 40Kb 777 candidates

VII/143/

T-f-v CRAMPTON D., SCHADE D., COWLEY A.P.: 1985, Astron. J. 90, 987 CRAMPTON D., COW-LEY A.P., SCHMIDTKE P.C.: 1988 , Astron. J. 96, 816 CRAMPTON D., COW-LEY A.P., HARTWICK F.D.A: 1990 , Astron. J. 100, 47

Id.: IAU α, δ (1950) m_{B}, z Remarks

Nearby Galaxies Catalogue 830Kb 2367 galaxies

VII/145/

T-f-v Tully R.B, Cambridge University Press (1988)

Id.: NGC, UGC, Anon α, δ (1950); ℓ, b; L, B Morph.Type, Diameters, i B_{T}^{b,i}, H_{T}^{b,i}, B − H, Abs. V_{h}, V_{0} W_{20}, W_{R}^{i}, W_{D}^{i} R, XY Z position L_{B}, M_{H}, M_{T} Groups, clusters by exchange from NASA-Goddard

Photometric Catalogue of Northern Bright Galaxies 108Kb 791 galaxies

VII/147/

B-N-T-f-v Kodaira K., Okamura S., Ichikawa S., Hamabe M., Watanabe M. , National Astronomical Observatory, Mitaka, Tokyo, 181 Japan (1992)

Id.: NGC, other α δ (1950), Morph. Type V, V_{total}, A_{V} Envelope Ellipse Eff. Surface Brightness Lum. Concentr. Index Notes by exchange from NAO-Japan

Positions of high-redshift luminous quasars from the HST Snap-Shot Survey 71Kb 528 quasars

VII/149/

N-T-f-v Schneider D.P., Bahcall J.N., Saxe D.H., Bahcall N.A., Doxsey R. Golombek D., Krist J., McMaster M., Meakes M., Lahav O. , Publ. Astron. Soc. Pac. 104, 678 (1992)

Id.: Name α, δ (J2000, B1950) Δα, Δδ, M_{V} , z, Finding chart frame number by exchange from NASA-Goddard by exchange from H.Andernach, IAC, Tenerife, Spain

Structure Parameters of Galactic Globular Clusters 68Kb 167 clusters

VII/151/

B-N-T-f-v Webbink R.F. , IAU Symp. 113 “Dynamics of star clusters”, Ed. J. Goodman & P. Hut, 541 (1985)

Id.: IAU, NGC/IC/other α, δ (1950) V_{HB}, E(B − V), (B − V)_{g}, log θ_{t}, log θ_{c}, V_{t}, σ_{c} \underline{Derived parameters}: ℓ, b R_{0}, (X, Y, Z, R_{GC}) [m/H], M_{v}, r_{c}, r_{t}, c_{der} log t_{r}, log ρ_{0}, σ_{0}, v_{esc} See also VII/195 Catalogue: VII/151 (6 tables)

Catalogue  References  Contents  Remarks

Third Reference Catalogue of Bright Galaxies (RC3) 8.2Mb 23022 galaxies

VII/155/

N-T-f-v de Vaucouleurs G., de Vaucouleurs A., Corwin H.G., Buta R.J., Paturel G., Fouqué P., Springer-Verlag, New-York (1991)

Id.: NGC/IC, UGC, ESO, MCG, CGCG, PGC, other

α, δ (1950, 2000) ℓ, b ; L, B Type, T, L R_{25}, D_{25}, A_{e}, D_{0} A_{g}, A_{i}, A_{21} extinctions B_{T} , m_{B}, m_{F IR}, B_{T}^{0} mag. Color Indices m_{21}, W_{20}, W_{50}, HI V_{21}, V_{opt}, V_{GSR}, V_{3K} Replaces VII/137 by exchange from NASA-Goddard

The Extended 12μm Galaxy Sample 104Kb 893 galaxies

VII/157/

N-T-f-v Rush B., Malkan M., Spinoglio L., Astrophys. J. Suppl. Ser. 89, 1 (1993)

Id.: α, δ (1950) IRAS fluxes z, reference Object type by exchange from NASA-Goddard

A Revised and Updated Catalog of Quasi-Stellar Objects 4.5Mb 7315 quasars

VII/158/

N-T-f-v Hewitt A., Burbidge G. , Astrophys. J. Suppl. Ser. 87, 451 (1993)

Id.: IAU, other α, δ (B1950, J2000) UBV, var. flag z(em.), z(abs.) References, Remarks Supersedes VII/117 and VII/128

CATALOGUE OF 7582 GALAXIES TOWARDS THE COMA SUPERCLUSTER OF GALAXIES 488Kb 7582 objects

VII/159/

N-T-f-v SLEZAK E., MARS G., BI-JAOUI A., BALKOWSKI C., FONTANELLI P.: 1988, Astron. Astrophys. Suppl. Ser. 74, 83 12-Feb-1994 Id.: Number α, δ (J2000) m_{B}, area in pixels Morph.Type

NEARBY GALAXIES 113Kb 276 galaxies

VII/161/

N-T-f-v SCHMIDT K.-H., BOLLER T., PRIEBE A.: 1993 , Astron. Nachr. 313, 189 (1992) Astron. Nachr. 314, 371 (1993)

Id.: NGC, IC, UGC, ESO, ... IRAS α, δ (1950); ℓ, b; L, B v_{⊙}, v_{LG}, Δv V_{max}, maxV_{max} m_{B}, M_{B}, B_{T}, B_{T} (0, i), type (B − V)_{T}^{0} , (U − B)_{T}^{0} log D_{25}, log R_{25}, log D_{0} S, ΔS, log F_{IR} log M_{HI}, log M_{dust} Θ, A_{g}, x_{2}

Compilation of published redshifts for ACO Galaxy Clusters 67Kb 1059 records

VII/165A/

N-T-f-v Andernach H. , ASP Conf. Ser. 15, 279 (1991)

Id.: Abell or ACO α, δ (1950) (center) Richness, distance and Bautz-Morgan class m_{10} (10th brightest) z, Abell radius log(z_{meas}/z_{est})

Catalogue  References  Contents  Remarks

Catalogue of Double Galaxies in the region of the SGP 67Kb 630 systems

VII/168/

N-T-f-v Zou Z., Chen J., Tang X., Bian Y., Bull. Inf. CDS 37, 173 (1989)

Id.: Seq α, δ (1950) Morph.Type, dimensions m_{B}, separation

OPTICAL IDENTI-FICATIONS OF IRAS POINT SOURCES 108Kb 759 sources

VII/169/

M-N-T-f-v WOLSTENCROFT R.D., et al.: 1986, Mon. Not. R. Astron. Soc.223, 279 LEGGETT S.K., et al.: 1987, Mon. Not. R. Astron. Soc.227, 563 WANG G., et al.: 1991, Mon. Not. R. Astron. Soc.248, 112 Id.: IRAS, counterpart α, δ(1950) (COSMOS or IRAS) offset, Morph.Type, B_{J}, notes F (far-IR), luminosity ratio of far-IR to COSMOS-B by exchange from H.Andernach, IAC, Tenerife, Spain

Catalogue of Southern Peculiar Galaxies and Associations 953Kb 6445 galaxies

VII/170/

N-T-f-v Arp H.C., Madore B.F., Cambridge University Press (1987)

Id.: AM

α, δ (1950) ; ℓ, b ; L, B (X, Y) on ESO/SERC plate

Diameter

Coded description

University of Michigan Catalog of emission-line Objects 27Kb 698 objects

VII/171/

N-T-f-v MacAlpine G.M., Smith S.B., Lewis D.W., Astrophys. J. Suppl. Ser. 34, 95 (1977); Astrophys. J. Suppl. Ser. 35, 197 (1977); Astrophys. J. Suppl. Ser. 35, 203 (1977); Astrophys. J. Suppl. Ser. 36, 587 (1978) MacAlpine G.M., Williams G.A. (1981), Astrophys. J. Suppl. Ser. 45, 113 MacAlpine G.M., Feldman F.R., Astrophys. J. 261, 412 (1982)

Id.: Number, other

α, δ (1950)

From

H.Andernach

FIRST BYURAKAN SUR-VEY (Markarian galaxies) 372Kb 1469 galaxies

VII/172/

N-T-f-v MARKARIAN B.E., LIPOVETSKY V.A., STEPANIAN J.A., ERAS-TOVA L.K., SHAPOVAL-OVA A.I., Comm. SAO (USSR) 62 (1989)

Id.: Mrk, KUG, Zw, NGC, IC, other α, δ (1950) z_{0}, dimensions UBV, IRAS fluxes Morph.Type Spectrum type

Supersedes VII/61A/

A CATALOGUE OF SEYFERT GALAXIES 417Kb 959 objects

VII/173/

N-T-f-v LIPOVETSKY V.A., NEIZVESTNY S.I., NEIZVESTNAYA O.M. , Comm. SAO (USSR) 55 (1988)

Id.: IAU-like, other

α, δ (1950) z, Seyfert Type m_{B}, M_{B}, E(B − V) Morph.Type, Dimensions UBV R photometry Fluxes: X-ray, Hβ, [OIII], JHKLN, IRAS, 6cm, 11cm References from H. Andernach

Catalogue  References  Contents  Remarks

A Catalogue of Morphological types in 55 rich clusters of galaxies 412Kb 5725 galaxies

VII/174/

N-T-f-v Dressler A., Astrophys. J. Suppl. Ser. 42, 565 Id.: Abell, number α, δ (1950) ; X, Y Morph.Type m_{vis}, m_{bulge} ε (ellipticity)

Edinburgh-Durham Southern Galaxy Catalogue - the cluster Catalog 71Kb 737 clusters

VII/175/

N-T-f-v LUMSDEN S.L., NICHOL R.C., COLLINS C.A., GUZZO L., Mon. Not. R. Astron. Soc. 258, 1 (1992)

Id.: EDCC, A, J-survey α, δ (1950) m_{1}, m_{3}, m_{10} Radius, N(gal), N(bg) N(locations) by exchange from H.Andernach, IAC, Tenerife, Spain

Photometric and spectroscopic data for elliptical galaxies 100Kb 479 galaxies

VII/176A/

N-T-f-v Faber S.M., Wegner G., Burstein D., Davies R.L., Dressler A., Lynden-Bell D., Terlevich R.J., Astrophys. J. Suppl. Ser. 69, 763 (1989)

Id.: Number, other ℓ, b, A_{B}, N(gr) v_{⊙}, v_{group}, (B − V)_{0}, B_{T}, SB_{e} Diameters: A_{e}, D_{n} Mg_{2} index by exchange from H.Andernach, IAC, Tenerife, Spain

A Compilation of redshifts and Velocity Dispersions for Abell Clusters 66Kb 774 clusters

VII/177/

N-T-f-v Struble M.F., Rood H.J., Astrophys. J. Suppl. Ser. 77, 363 (1991)

Id.: Abell α, δ (1950) z(adopted), N_{obs}, σ z(SR1), N_{obs}, σ z(new), N_{obs}, σ Notes and references by exchange from H.Andernach, IAC, Tenerife, Spain

* See also VII/95 * First archived as J/ApJS/77/363/

An Optical Catalog of Extragalactic Emission-Line Objects similar to Quasi-Stellar Objects 154Kb 935 galaxies

VII/178/

N-T-f-v Hewitt A., Burbidge G. , Astrophys. J. Suppl. Ser. 75, 297 (1991)

Id.: IAU, other α, δ (1950) m, z_{em} References by exchange from H.Andernach, IAC, Tenerife, Spain

* First archived in J/ApJS/75/297/

Kiso Survey for Ultraviolet-Excess galaxies. XVIII 1.1Mb 8162 galaxies

VII/179A/

T-f-v Takase B., Miyauchi-Isobe N., Publ. Nat. Astron. Obs. Japan: 3, 169 (1993)

Id.: KUG, other α, δ (1950) Morphological type Apparent size, m UV excess \underline{Note}: Those objects which appeared repeatedly in the previous version have been omitted. by exchange from NAO-Japan supersedes VII/150 Erratum taken into account

Catalogue  References  Contents  Remarks

Catalogue of galaxies in the central 3.5 degs of the Fornax Cluster, and in five nearby groups 366Kb 3466 galaxies

VII/180/

N-T-f-v Ferguson H.C., Astron. J. 98, 367 (1989) Ferguson H.C., Sandage A., Astron. J. 100, 1 (1990)

Id.: Number, Name, (IAU), (plate) α, δ (1950) Membership status Morph.Type B_{T}, r_{e}, D, sources V_{⊙} + σ by exchange from H.Andernach, IAC, Tenerife, Spain Supersedes VII/160/ and J/AJ/100/1/

A CATALOG OF X-ray MEA-SUREMENTS OF SE-LECTED SAMPLES OF AC-TIVE GALAXIES AND NU-CLEI 407Kb 1176 objects

VII/181/

N-T-f-v

DELLA  CECA  R.,

PALUMBO G.G.C., PER-SIC M., BOLDT E.A., DE ZOTTI G., MARSHALL E.E.: 1990, Astrophys. J. Suppl. Ser. 72, 471 Id.: PG, other α, δ (1950) z, m_{B} F (X), E, Observation time

from H. Andernach

A Wide-Field Multicolor Survey for High-Redshift Quasars, z ≥ 2.2 58Kb 169 quasars

VII/182/

N-T-f-v Warren S.J., Hewett P.C., Irwin M.J., Osmer P.S., Astrophys. J. Suppl. Ser. 76, 1 (1991) Warren S.J., Hewett P.C., Osmer P.S., Astrophys. J. Suppl. Ser. 76, 23 (1991)

Id.: IAU-style α, δ (1950) Obs. date, tel., z u, b_{j}, v, or, r, i magnitudes W (Lyα), W (CIV) α-spectral index by exchange from NASA-Goddard by exchange from H.Andernach, IAC, Tenerife, Spain Supersedes VII/148/, J/ApJS/76/1/ and J/ApJS/76/23/

The Cluster System of the Large Magellanic Cloud 85Kb 1762 clusters

VII/183/

N-T-f-v Kontizas M., Morgan D.H., Hatzidimitriou D., Kontizas E., Astron. Astrophys. Suppl. Ser. 84, 527 (1990)

Id.: Seq, NGC, Hodge, ... α, δ (1950) Cluster diameter Class (compact or loose)

Survey of 108 E-S0 RSA Galaxies 3.7Mb 108 galaxies

VII/184/

N-T-f-v Michard R., Marchal J., Astron. Astrophys. Suppl. Ser. 98, 29 (1993) Astron. Astrophys. 283, 779 (1994) Astron. Astrophys. Suppl. Ser. 105, 481 (1994)

Id.: NGC Morphology B_{T}^{0} , V_{T}, M_{T} (B) μ_{e}^{∗}(V), μ_{e}^{∗}(B) log(D_{25}), log(R_{25}), log(R_{e}^{∗})

Isophotal brighnesses, distances and coefficients according to Carter’s principles

IRAS 1.2 Jy IRAS Redshift Survey 998Kb 9897 sources

VII/185/

N-T-f-v Strauss M., Davis M., Yahil A., Huchra J.P., Astrophys. J. 361, 49 (1990) Strauss M., Huchra J.P., Davis M., Yahil A., Fisher K.B., Tonry J. , Astrophys. J. Suppl. Ser. 83, 29 (1992) Fisher K.B., Huchra J.P., Strauss M.A., Davis M., Yahil A., Schlegel D., Astrophys. J. Suppl. Ser. 100, 69 (1995)

Id.: IRAS α, δ (1950) IRAS fluxes Optical counterpart z by exchange from NASA-Goddard

Catalogue  References  Contents  Remarks

CATALOGUE OF GALAX-IES IN A3744 7.5Kb 51 galaxies

VII/186/

N-T-f-v CAMERON R.A., PhD Thesis, Sydney (1988)

Id.: Seq. α, δ (1950) V_{r}(em), V_{r}(abs)

Catalog of 706 pulsars 550Kb 706 pulsars

VII/189/

f-v Taylor J.H., Manchester R.N., Lyne A.G., Camilo F. , http://pulsar.princeton.edu/ Id.: PSR (J and B) α, δ (B1950, J2000) with σ ℓ, b ; μ_{α}, μ_{δ} P, Ṗ with σ, Epoch Dispersion measure Pulse profile F (400, 1400) MHz Orbital elements for binary pulsars Derived parameters: d, z log L log τ log Ė log B References

Supersedes VII/156

ZWICKY GALAXY CATA-LOG 2.7Mb 28840 galaxies

VII/190/

T-f-v ZWICKY F., et al.: 1961-1968, Calif. Inst. of Techn., Pasadena, 6 vols. Kalinkov M., Valtchanov I., Kuneva I. , Institute of Astronomy, Bulgarian Academy of Sciences Id.: ZNr, Z, Other α, δ (1950) Notes Additional magnitudes & Velocities Supersedes the galaxy part in VII/49

A catalogue of southern dark clouds 68Kb 1101 nebulae

VII/191/

M-N-f-v Hartley M., Manchester R.N., Smith R.M., Tritton S.B., Goss W.M., Astron. Astrophys. Suppl. Ser. 63, 27 (1986)

Id.: DCld α, δ (1950) Dimensions Density class Plate, Comments δ < −33^{∘} Prepared at CDS See also VII/7/

Arp’s Peculiar Galaxies 75Kb 592 galaxies

VII/192/

T-f-v Webb D., Sky & Telescope, 92-1, 92 (1996)

Id.: Arp, other α, δ (2000) Size, Orientation Focal length of camera by exchange from NASA-Goddard See also VII/74/

The CfA Redshift Survey, Version June 1995 5.3Mb 58738 redshifts

VII/193/

N-T-f-v

Huchra  J.P.,  Geller

M.J., Clemens C.M., Tokarz S.P., Michel A., Harvard-Smithsonian Center for Astrophysics (1995)

Id.: NGC/IC, UGC, ESO, other α, δ (1950) m_{B}, B_{T} V_{r}, source, σ Morphology supersedes VII/164 and J/ApJS/99/391

Photometry of Shakhbazyan Groups 774Kb 373 groups

VII/196/

T-f-v Stoll D., Tiersch H., Braun M., Cordis L., Astron. Nachr. 314 225, 317 (1993); Astron. Nachr. 315 11, 97 (1994); Astron. Nachr. 317, 239, 315, 383 (1996); Astron. Nachr. 318, 7, 89, 149 (1997)

Id.: Shk group number \underline{For each galaxy in group}:

α, δ (2000) m_{V} or m_{R} V_{r} or z Notes

Mark III Catalog of Galaxy Peculiar Velocities: Distances for Groups and Single Galaxies 1.6Mb 3012 galaxies

VII/198/

T-f-v Willick J.A., Courteau S., Faber S.M., Burstein D., Dekel A., Kolatt T., Strauss M.A., Astrophys. J. 446, 12 (1995); Astrophys. J. 457, 460 (1996); Astrophys. J. Suppl. Ser. 109, 333 Id.: Mark3, RC3, UGC, ESO, other α, δ (1950) ; ℓ, b Dimensions, Morphology cz, Hi flux Groups of Galaxies by exchange from NASA-Goddard

128  Catalogues available at CDS  ( January 19, 2008)

Catalogue  References  Contents  Remarks

Data Base of Surface Photometries of the Milky Way 4.0Mb 4 FITS images

VII/199/

T-f Wolfgang Hovest , Astron. Astrophys. Suppl. Ser. 128, 417 (1998)

4 FITS images

Redshift Survey at the NGP 36Kb 320 galaxies

VII/200/

T-f-v Willmer C.N.A. , Koo D.C. , Ellman N. , Szalay A.S., Kurtz M.J., Astrophys. J. Suppl. Ser. 104, 199 + Erratum 107, 823 (1996)

Id.: Running α, δ (2000) z, b_{J}, b_{J} − r_{F}

Zwicky Clusters of Galaxies Catalog 257Kb 9133 clusters

VII/201/

T-f-v Kalinkov M., Valtchanov I., Kuneva I., Bull. Inf. CDS 49 (unpublished)

Id.: CGCG number + Fields α, δ (1950) Cluster type, population, Diameter

A catalog of parameters for globular clusters on the Milky Way 84Kb 147 clusters

VII/202/

T-f-v Harris W.E., Astron. J. 112, 1487 (1996)

Id.: NGC, other α, δ (2000) ; ℓ, b R_{⊙}, R_{GC}, X, Y, Z distances E(B − V), integrated magnitudes & colors Integrated Spectral type, HB types [Fe/H], V_{r}, core descriptors See also VII/151 Supersedes VII/195

The Las Campanas Redshift Survey 7.9Mb 82360 galaxies

VII/203/

T-f-v Shectman S.A., Landy S.D., Oemler A., Tucker D.L., Lin H., Kirshner R.P., Schechter P.L. , Astrophys. J. 470, 172 (1996)

Id.: LCRS α, δ (1950) Isophotal m cz, σ List of fields From AAS CD-ROM #7 700sq.^{∘} near SGP

Galaxy properties at NGP 2.1Mb 36402 galaxies

VII/204/

T-f-v Odewahn S.C., Aldering G., Astron. J. 110, 2009 (1995)

Id.: POSS-I+No m_{O}, E − O C_{31} concentration index

Photometric Magnitudes and Colors of Galaxies 7.8Mb 7288 galaxies

VII/206/

T-f-v Prugniel P., Heraudeau P., Astron. Astrophys. Suppl. Ser. 128, 299 (1998)

Id.: NGC, IC, UGC, PGC α, δ (2000) B_{t}, (B−V)_{e}, (U−B)_{e}, (V − R)_{e}, (R − I)_{e} T_{phot}, Quality Aperture photometry Bibliography Supersedes VII/167

Catalogue  References  Contents  Remarks

Faint Galaxies at the North Galactic Pole: The Catalogue 6.7Mb 60852 galaxies

VII/208/

T-f-v Infante L., Pritchet C.J., Hertling G., J. Astron. Data 1, N^{∘} 2 (1995)

Id.: NGPFG α, δ (1950) J and F magnitudes S_{xx}S_{yy}S_{sy} cov. Area profile parameters From JAD#1 CD-ROM

Catalog of Galaxies behind the Milky Way: vols 1-5 (ℓ = −7^{∘} to ℓ = 68^{∘}, ℓ = 210^{∘} to ℓ = 250^{∘}) 3.0Mb 19530 galaxies

VII/209A/

T-f-v Saito M., et al.; Roman A.T., et al., Publ. Astron. Soc. Japan 42, 603 (1990) ; 43, 449 (1991); 48, 679 (1996) ; 50, 37 (1998) ; Astrophys. J. Suppl. Ser. 127, 27 (2000)

Id.: CGMW, ESO/SERC, other α, δ (1950), (ℓ, b) Major/Minor diameters Features Cross-identifications Supersedes VII/140 by exchange from NAO-Japan

Hickson’s Compact groups of Galaxies 94Kb 100 groups

VII/213/

T-f-v Hickson P., Astrophys. J. 255 (1982) Hickson P., Kindl E., Auman J.R., Astrophys. J. Suppl. Ser. 70, 687 (1989) Mendes de Oliveira C., Hickson P., Astrophys. J. 380, 30 (1991) Hickson P., Mendes de Oliveira C., Huchra J.P., Palumbo G.G.C., Astrophys. J. 399, 353 (1992) Mendes de Oliveira C., Hickson P., Astrophys. J. 427, 684 (1994)

\underline{Groups}: Id.: Number, other α, δ (1950) Group classification N(galaxies) Angular diameter Total magnitude Magnitude of brightest Velocity Dispersion \underline{Galaxies}: z, photometry, Morphology From H. Andernach

Supersedes VII/85/ See also J/AJ/123/2976

APS Galaxies in the North Galactic Pole 66Mb 217768 galaxies

VII/214/

T-f-v

Id.: APS (POSS/MLN + number) α, δ (B1950); x, y ϵ, dimensions C_{31}, C_{32}, C_{21} E(B−V), μ_{E}, μ_{O}

A catalogue of H-alpha emission regions in the Southern Milky Way 25Kb 182 nebulae

VII/216/

T-f-v Rodgers A.W., Campbell C.T., Whiteoak J.B., Mon. Not. R. Astron. Soc. 121, 103 (1960)

Id.: RCW

α, δ (B1950), ℓ, b

Dimension

Brightness estimate

Cross-

identifications

Catalogue of southern stars embedded in nebulosity 24Kb 136 nebulae

VII/218/

T-f-v van den Bergh S., Herbst W., Astron. J. 80, 208 (1975)

Id.: VDBH α, δ (1950) Type of nebula Diameter Color, Absorption Remarks See also VII/21

Revised Flat Galaxy Catalogue 617Kb 4444 galaxies

VII/219/

T-f-v Karachentsev I.D., Karachentseva V.E., Kudrya Y.N., Sharina M.E., Parnovsky S.L. , Bull. Special Astrophys. Obs. 47, 5 (1999)

Id.: RFGC, FGC, PGC α, δ (J2000, J1950) (a, b)_{O}, (a, b)_{E} B_{T}, Notes Supersedes VII/162

Catalogues available at CDS  ( January 19, 2008)  131

Catalogue  References  Contents  Remarks

Barnard’s Catalogue of 349 Dark Objects in the Sky 56Kb 349 nebulae

VII/220A/

T-f-v Barnard E.E., Carnegie Institution of Washington (1927)

Id.: Barnard α, δ (1875, 2000)

The PSCz catalogue 27Mb 18351 sources

VII/221/

T-f-v Saunders W., Sutherland W.J., Maddox S.J., Keeble O., Oliver S.J., Rowan-Robinson M., McMahon R.G., Efstathiou G.P., Tadros H., White S.D.M., Frenk C.S., Carraminana A., Hawkins M.R.S., Mon. Not. R. Astron. Soc. 317, 55 (2000)

Id.: IRAS, PGC

α, δ (B1950) (IR + optical) IRAS data (fluxes etc) cz, distance 100μm background Qualities FORTRAN application routines

See II/125

The Canada-France Redshift Survey 105Kb 1010 sources

VII/225A/

f-v

[Data in Simbad]

Lilly S.J., Le Fèvre O., Crampton D., Hammer F., Tresse L. , Astrophys. J. 455, 50; 455, 60; 455, 75; 455, 88; 455, 96 Id.: CFRS α, δ (J2000) I_{AB}, (V − I)_{AB}, Q z, confidence class

A Catalogue of Galactic Supernova Remnants, 2001 December version 24Kb 231 SNRs

VII/227/

f-v Green D.A., Proc. IAU Colloq. 145 ”Supernovae and Supernovae Remnants”, Cambridge Univ. Press, p. 419 (1996); 2001 Version Id.: G-based α, δ (1950) ; ℓ, b Type, Angular size F (1GHz) by exchange from NASA-Goddard Supersedes VII/211

Las Campanas Infrared Survey 21Mb 59284 sources

VII/228/

f-v The Las Campanas Infrared Survey. , Astrophys. J. 560, L131 (2001) Mon. Not. R. Astron. Soc. 332, 617 (2002) Astrophys. J. 570, 54 (2002)

Id.: Tile+Number

α, δ (J2000)

V RIz′H at 2′′ and 4′′ CDFS and HDFS fields

Catalog of bright diffuse Galactic nebulae 205Kb 330 nebulae

VII/231/

f-v Cederblad S., Lund Medd. Astron. Obs. Ser. II, 119, 1 (1946)

Id.: Ced, other α, δ (B1900) Dimensions, Spectrum log a_{1}, Distance Associated star Notes and bibliography

The nebular complexes of the Magellanic Clouds. 51Kb 524 nebulae

VII/232/

f-v Davies R.D., Elliot K.H., Meaburn J. , Mem. R. Astron. Soc. 81, 89 (1976)

Id.: DEM, Henize

α, δ (B1950)

Cloud Dimensions Intensity Scale Comments

Catalogue  References  Contents  Remarks

DENIS I-band catalog of Galaxies 3.1Mb 42258 galaxies

VII/234/

f-v

Paturel  G.,  Petit

C., Rousseau J., Vauglin I., Astron. Astrophys. 405, 1 (2003)

Id.: PGC, other α, δ (J2000) Dimensions Da/bθ Supersedes J/A+AS/135/133

Catalog of Interacting Galaxies 344Kb 3160 records

VII/236/

f-v Vorontsov-Velyaminov B.A., Noskova R.I., Arkhipova V.P. , Astron. & Astrophys. Trans. 20, 717 (2001)

Id.: VV, NGC, MCG, Zw, Mrk, Arp, UGC, other α, δ (J2000) m_{B}, V_{r}, z, Morph. type

HYPERLEDA Catalog of Galaxies, Version 2003 81Mb 983261 galaxies

VII/237/

f-v Paturel G., Petit C., Prugniel P., Theureau G., Rousseau J., Brouty M., Dubois P., Cambresy L., Astron. Astrophys. 412, 45 Id.: PGC, other α, δ (J2000) Morphological type D_{25}, R_{25}

Supersedes VII/119

HYPERLEDA II: Homogenized HI data 1.8Mb 16781 galaxies

VII/238/

f-v Paturel G., Theureau G., Bottinelli L., Gouguenheim, Coudreau-Durand N., Hallet N., Petit C., Astron. Astrophys. 412, 57 Id.: PGC, other α, δ (J2000) log(2V_{M}^{sin i}), m_{21} + σ V_{H I} + σ Notes

Puzzle Solutions in the NGC/IC 6.5Mb 15317 sources

VII/239A/

f-v Harold G. Corwin Jr , The NGC/IC Project Id.: NGC / IC Comments bout some of the sources

Millennium Galaxy Catalogue

58Mb 69594 galaxies

VII/240/

f-v Liske J., Lemon D.J., Driver S.P., Cross N.J.G., Couch W.J. , Mon. Not. R. Astron. Soc. 344, 307 (2003) Driver S.P., Liske J., Cross N.J.G., De Propris R., Allen P.D. , Mon. Not. R. Astron. Soc. 360, 81 (2005)

Id.: MGC, SDSS, NED

α, δ (J2000); x, y Classification Morphology Dimensions Surface brightness Photometries from SDSS, 2dF, 2MASS, NED Redshifts

The 2dF QSO Redshift Survey (2QZ) 16Mb 49425 quasars

VII/241/

f-v Croom S.M., Smith R.J., Boyle B.J., Shanks T., Loaring N.S., Miller L., Lewis I.J., Mon. Not. R. Astron. Soc. 349, 1397 (2004)

Id.: 2QZ α, δ (J2000) b_{J}, u − b_{J}, b_{J} − r z at 2 dates Quality, S/N F (21cm) F_{x} Supersedes VII/223

Catalogue  References  Contents  Remarks

A catalog of LEDA galaxies with DENIS measurements 264Mb 753153 galaxies

VII/242/

f-v Paturel G., Vauglin I., Petit C., Borsenberger J., Epchtein N., Fouqué P., Mamon G., Astron. Astrophys. 430, 751 (2005)

Id.: PGC, other

α, δ (J2000) in (IJK)

log D_{25}, log R_{25} (opt., IJK) IJK total + Kron magnitudes IJK Object types

The Sloan Digital Sky Survey Quasar Catalog III. Third Data Release 19Mb 46420 QSOs

VII/243/

f-v Schneider D.P., Hall P.B., Richards G.T., et al. , Astron. J. 130, 367 (2005)

Id.: SDSS ; α, δ (J2000) ; z

See also III/241

Atlas and Catalog of Dark Clouds 540Kb 5289 clouds

VII/244A/

f-v Dobashi K., Uehara H., Kandori R., Sakurai T., Kaiden M., Umemoto T., Sato F. , Publ. Astron. Soc. Japan 57, S1 (2005)

Id.: Seq

ℓ, b (galactic) Extent, Area, A_{V} Flags and Notes

Quasars and Active Galactic Nuclei (12th Ed.) 15Mb 108080 qso+agn

VII/248/

f-v Véron-Cetty M.P., Véron P., Astron. Astrophys. 455, 773 (2006)

Id.: most common names α, δ (J2000, B1950) UBV z, remark F (6cm) , F (20cm) Classification References

Supersedes VII/235

Second Data Release of the 6dF Galaxy Survey 36Mb records

VII/249/

f-v Jones D.H., Saunders W., Read M., Colless M., Publications - Astron. Soc. Australia, 22, 277 (2005)

Id.: ; α, δ (J2000) ; σ

The 2dF Galaxy Redshift Survey (2dFGRS) 207Mb 245591 sources

VII/250/

f-v Colless M.M., et al. , Mon. Not. R. Astron. Soc. 328, 1039 (2001)

Id.: 2dF α, δ (J2000) b_{J} z, S/N, quality Supersedes VII/226

VIII. Radio and Far-IR Data

Catalogue  References  Contents  Remarks

The Third Cambridge Catalogues (3C and 3CR) 71Kb 328 sources

VIII/1A/

f-v Edge D.O., Shakeshaft J.R., McAdam W.B., Baldwin J.E., Archer S. , Mem. Roy. Astron. Soc. 68, 37 (1959) Bennett A.S., Mem. Roy. Astron. Soc. 68, 163 (1961)

Names: 3C, 3CR

α, δ (1950)

Flux at 159 and 178MHz Diameters Notes and Remarks

Optical Catalog of Radio Galaxies 185Kb 495 sources

VIII/3/

T-f-v Burbidge G.R., Crowne A.H., Astrophys. J. Suppl. Ser. 40, 3 (1979)

Id.: + Other names

α, δ (1950)

Finder chart reference Optical type code Radio description code m_{V} , B − V, U − B z Spectral line type F (Radio)

By exchange from NASA-Goddard Previously under number VII/33/

Fourth Cambridge Radio Survey Catalogue (4C) 307Kb 4844 sources

VIII/4/

N-T-f-v PILKINGTON J.D.H., SCOTT P.F.: 1965, Mem. Roy. Astron. Soc. 69, 183

GOWER J.F.R., SCOTT P.F., WILLS D.: 1967 , Mem. Roy. Astron. Soc. 71, 49

Id.: 4C α, δ (1950) ℓ^{II}, b^{II} F (178MHz) Error class Remarks

By exchange from NASA-Goddard Previously under number VII/53/

Extragalactic Radio Sources having Flux Densities > 1 Jy at 5 GHz 415Kb 518 sources

VIII/5/

T-f-v KUHR H., WITZEL A., PAULINYTOTH I.I.K., NAUBER U.: 1981, Astron. Astrophys. Suppl. 45, 367 Id.: 3C, Source name α, δ (1950), σ reference m_{v}, z spectral index Flux density: ν, F_{ν}, σ_{F} Previously under number VII/78/

1400  MHz  Sky

Survey confusion-limited maps covering −5^{∘} < δ < +85^{∘} 42Mb 144 FITS files

VIII/6/

B-I-M-N-T-f-v Condon J.J., Broderick J.J., Astron. J., 90, 2540 (1985) Astron. J., 91, 1051 (1986)

Observation Date

α, δ (1950) of map center 144 FITS files (2 × 2′ pixel size) of calibrated flux By exchange from NASA-Goddard Previously under number VII/97

The Hat Creek High-Latitude HI Survey 127Mb 134932 records

VIII/7A/

B-N-T-f-v Heiles C., Habing H.J.: 1974 , Astron. Astrophys. Suppl. Ser. 14, 557 Heiles C., Jenkins E.B.: 1976 , Astron. Astrophys. 46, 333 α, δ (1950) ; ℓ, b Obs. date, status Antenna temperatures in 100 filters from -90.76 to +71.76 km/s |b| > 10^{∘}, δ > −30^{∘} FITS maps by exchange from NASA-Goddard

Catalogue  References  Contents  Remarks

The Parkes high-latitude HI Survey 9.6Mb 9891 records

VIII/8A/

B-N-T-f-v Cleary M.N., Heiles C., Haslam C.G.T., Astron. Astrophys. Suppl. Ser. 36, 95 (1979) Heiles C., Cleary M.N. , Australian J. Phys. Suppl.47 (1979) α, δ (1950) ; ℓ, b V_{LSR} T_{A} in 64 channels |b| > 10^{∘}, δ < −30^{∘} FITS maps by exchange from NASA-Goddard

The Argentina High-Latitude HI Survey 21Mb 55470 profiles

VIII/9/

N-f-v Colomb F.R., Poeppel W.G.L., Heiles C. , Astron. Astrophys. Suppl. Ser. 40, 47 (1980) α, δ (1950) ; ℓ, b Obs. date and time T_{A} for 56 channels |b| > 10^{∘}, δ < −25^{∘} ΔV_{r} = 2km/s, beam=30′ by exchange from NASA-Goddard

The Berkeley low-latitude Survey of Neutral Hydrogen 45Mb 38961 profiles

VIII/11/

N-f-v Weaver H., Williams D.R.W., Astron. Astrophys. Suppl. Ser. 8, 1 (1973)

T_{A} for 239 points (Δν=10kHz) 10^{∘} ≤ ℓ ≤ 250^{∘} −10^{∘} ≤ b ≤ +10^{∘} Δℓ = 0^{∘}.5 Δb = 0^{∘}.25 by exchange from NASA-Goddard

ORION A EMISSION-LINE SURVEY FROM 330.5 TO 360.1 GHz 5.0Kb 75 spectra

VIII/12/

N-f JEWELL P.R., HOLLIS J.M., LOVAS F.J., SNYDER L.E.: 1989 , Astrophys. J. Suppl. Ser. 70, 833 Figures 1 to 39 (75 spectra) as FITS tables by exchange from NASA-Goddard

A New Catalog of 53522 4.85GHz sources 2.2Mb 53522 sources

VIII/13/

B-M-N-T-f-v Becker R.H., White R.L., Edwards A.L. , Astrophys. J. Suppl. Ser. 75, 1-229 (1991)

Id.: IAU α, δ (1950) F (4.85GHz) Extended source flag 4.85/0.365 GHz index by exchange from NASA-Goddard

The 87GB Catalog of Radio Sources covering 0 to +75^{∘} at 4.85 GHz 8.6Mb 54579 sources

VIII/14/

B-N-T-f-v Gregory P.C., Condon J.J., Astrophys. J. Suppl. Ser. 75, 1011-1291 (1991) α, δ (B1950, J2000) ℓ, b F (4.85GHz) + σ FWHM axises Major axis PA Local sky level Map pixel position by exchange from NASA-Goddard Supreseded by VIII/40

PARKES Catalogue 1990 (PKSCAT90) 1.4Mb 8264 radio sources

VIII/15/

B-N-T-f-v Wright A., Otrupcek R.: 1990 , Australia Telescope National Facility ; Bull. Inf. CDS 41, 47 Id.: IAU, other α, δ (B1950, J2000), σ Mag., Class, z Fluxes for 80, 178, 408, 635, 1410, 2700, 5000, 8400, 22000 MHz References

Molonglo Reference Catalog of Radio Sources (MRC) 1.1Mb 12141 sources

VIII/16/

N-T-f-v Large M.I., Cram L.E., Burgess A.M., The Observatory 111, 72 (1991)

Source Name α, δ (1950, J2000) ; ℓ, b F (408MHz), morphology Cross-references by exchange from NASA-Goddard Supersedes VIII/2

Catalogue  References  Contents  Remarks

The 1.4GHz Northern Sky Catalog 2.2Mb 31524 sources

VIII/17/

N-T-f-v White R.L., Becker R.H., Astrophys. J. Suppl. Ser. 79, 331 (1992)

Id.: IAU name α, δ (1950) F (1.4, 4.85, 0.365 GHz) Spectral indexes 6/20/80cm −5^{∘} to +82^{∘} by exchange from NASA-Goddard

THE 6C SURVEY OF RA-DIO SOURCES - I. Zone δ > 80^{∘} 82Kb 1760 sources

VIII/18/

N-T-f-v BALDWIN J.E., et al.: 1985, Mon. Not. R. Astron. Soc. 217, 717 α, δ (1950) F (151MHz) (peak, int.) Contour map panel by exchange from NASA-Goddard See also VIII/21, 22, 23

Catalogues from a deep 327 MHz Westerbork Survey 1.4Mb 5850 sources

VIII/19/

N-T-f-v Wieringa M.H., Thesis, Leiden Univ. (1991) Bull. Inf. CDS 43, 17 (1993)

Id.: Westerbork name (64W2, 69W, 70W, 75W, 76W) α, δ (B1950) + σ F (327 or 608 MHz) + σ Source structure code Axises of fitted gaussians, PA S/N ratio Primary beam atten. Statistical weight

A Catalogue of 5 GHz Galactic Plane Sources 64Kb 915 sources

VIII/20/

T-f-v Haynes R.F., Caswell J.L., Simons L.W.J., Aust. J. Phys. Astrophys. Suppl 48, 1 (1979) ℓ, b ; α, δ (1950) T, F, Extents Comments −2^{∘} ≤ b ≤ 2^{∘} 190^{∘} ≤ ℓ ≤ 40^{∘} Keypunched at CDS

THE 6C SURVEY OF RADIO SOURCES - II. THE ZONE 30 < δ < 51, 08h30m < α < 17h30m 428Kb 8277 sources

VIII/21/

N-T-f-v HALES S.E.G., BALDWIN J.E., WARNER P.J.: 1988, Mon. Not. R. Astron. Soc. 234, 919 α, δ (1950) F (151MHz) (peak, int.) Contour map panel by exchange from NASA-Goddard See also VIII/18, VIII/22,23,24,25

The 6C survey of radio sources

- III. The zone 48 < δ < 68, 05h25 < α < 18h17 435Kb 8748 sources

VIII/22/

N-T-f-v HALES S.E.G., MASSON C.R., WARNER P.J., BALDWIN J.E.: 1990, Mon. Not. R. Astron. Soc.

246. 256 α, δ (1950) F (151MHz) (peak, int.) Contour map panel by exchange from NASA-Goddard See also VIII/18, VIII/21,23,24,25

The 6C survey of radio sources

- IV. The zone 67 < δ < 82, 0h < α < 24h 286Kb 5420 sources

VIII/23/

N-T-f-v HALES S.E.G., MAYER C.J., WARNER P.J., BALDWIN J.E.: 1991, Mon. Not. R. Astron. Soc. 251, 46 α, δ (1950) F (151MHz) (peak, int.) Contour map panel by exchange from NASA-Goddard See also VIII/18, VIII/21,22,24,25

Catalogue  References  Contents  Remarks

The 6C Survey of radio sources

- V. The zones 6C-Va (48 < δ < 68, 01h34 < α < 06h14) and 6C-Vb (48 < δ < 68, 17h16 < α < 20h24) 188Kb 3458 sources

VIII/24/

N-T-f-v HALES S.E.G., MASSON C.R., WARNER P.J., BALDWIN J.E., GREEN D.A.: 1993, Mon. Not. R. Astron. Soc. 262, 1057 α, δ (1950) F (151MHz) (peak, int.) Contour map panel by exchange from NASA-Goddard See also VIII/18, VIII/21,22,23,25

THE 6C SURVEY OF RADIO SOURCES - VI. THE CON-TINUOUS ZONE 6C-VI 30 < δ < 51, 22h < α < 09h05 341Kb 6752 sources

VIII/25/

N-T-f-v HALES S.E.G., BALDWIN J.E., WARNER P.J.: 1993, Mon. Not. R. Astron. Soc. 263, 25 α, δ (1950) F (151MHz) (peak, int.) Contour map panel by exchange from NASA-Goddard See also VIII/18, VIII/21,22,23,24

RESULTS of a FIVE-YEAR PROGRAM of MULTIFRE-QUENCY MONITORING of LOW-FREQUENCY VARI-ABLE RADIO SOURCES 103Kb 34 sources

VIII/26/

N-T-f MITCHELL K.J., DENNISON B., CONDON J.J., ALTSCHULER D.R., PAYNE H.E., O’DELL S.L., , BRODERICK J.J., Astrophys. J. Suppl. Ser. 93, 411 (1994)

Id.: IAU name S_{ν}, σ(S_{ν}) Observation date 318, 430, 606, 880, 1400MHz by exchange from NASA-Goddard

The Bell Laboratories H I Survey 79Mb 501 FITS files

VIII/28/

f-v Stark A.A., Gammie C.F., Wilson R.W., Bally J., Linke R.A., Astrophys. J. Suppl. Ser. 79, 77 (1992)

3×145 FITS files representing the sky in 145 velocity channels from −360km/s to +360km/s 66 FITS files representing T_{A} versus (αV_{r}) at donstant δ

A 1400-MHz Survey of 1478 Abell Clusters of Galaxies 143Kb 957 sources

VIII/29/

N-T-f OWEN F.N, WHITE R.A., HILLDRUP K.C., HANISCH R.J.: 1982, Astron. J. 87, 1083 Id.: Abell α, δ (1950) ; z Cluster richness, distance S(1400MHz), axises Distance from cluster center by exchange from NASA-Goddard

The 5C14, 5C15, and 5C16 RADIO SOURCE SURVEYS of the 5C12 AREA 51Kb 697 sources

VIII/30/

N-T-f-v Benn C.R., Kenderdine S., Mon. Not. R. Astron. Soc. 251, 253 (1991)

Id.: 5C12 α, δ (B1950) S(408MHz), S(1407MHz) α (spectral index) by exchange from NASA-Goddard

Revised source list for the Rees 38-MHz survey 334Kb 5859 sources

VIII/31/

N-T-f-v Hales S.E.G., Waldram E.M., Rees N., Warner P.J., Mon. Not. R. Astron. Soc. 274, 447 (1995)

Id.: 8C

α, δ (1950), size S_{p}, S_{int}, S/N by exchange from NASA-Goddard

Catalogue  References  Contents  Remarks
Observations  of  ex-

tragalactic radio sources with the Cambridge 5-km telescope at 5GHz 47Kb 104 sources

VIII/32/

f-v Jenkins C.J., Pooley G.G., Riley J.M. , Mem. R. Astron. Soc., 84, 61 (1977)

Id.: 3C, 4C α, δ (B1950) S(5GHz) + σ z Source extent Notes and References

A study of spectral characteristics of the RATAN RC-Catalog and preliminary optical identifications 185Kb 840 sources

VIII/33/

N-T-f-v Bursov N.N., Gol’neva N.E., Lipovka N.M., Soboleva N.S., Temirova A.V., Soobshch. Spets. Astrof. Obs. 63, 50 (1989)

Id.: RC α, δ (1950) ν, S(ν), Ref. Spectral index α Prepared at CDS cServer radio

The Parkes Half-Jansky Flat-Spectrum Sample 70Kb 323 sources

VIII/34/

T-f-v Drinkwater M.J., Webster R.L., Francis P.J., Condon J.J., Ellison S.L., Jauncey D.L., Lovell J., Peterson B.A., Savage A. , Mon. Not. R. Astron. Soc. 284, 85 (1997)

Id.: PKSCAT90 α, δ (1950) S(2.7GHz, 5.0GHz) α (sp.index) Optical counterpart

Radio Sources Observed with the Culgoora Circular Array 191Kb 2213 sources

VIII/35/

N-f-v

Slee  O.B.,  Aus-

tralian J. Phys. 48, 143-186 (1995)

Id.: IAU

α, δ (B1950) ; ℓ, b

S(80MHz), S(160MHz) Source dimension by exchange from H. Andernach

Second Bologna Survey 377Kb 10223 sources

VIII/36/

T-f-v Colla, Fanti et al, , Astron. Astrophys. Suppl. Ser. 1, 281 (1970) ; Astron. Astrophys. Suppl. Ser. 7, 1 (1972) ; Astron. Astrophys. Suppl. Ser. 11, 291 (1973); Astron. Astrophys. Suppl. Ser. 18, 147 (1974)

Id.: B2 ; α, δ (1950) ; σ

Third Bologna Survey 514Kb 13354 sources

VIII/37/

T-f-v

Ficarra  A.,  Grueff

G., Tomassetti G., Astron. Astrophys. Suppl. Ser. 59, 255 (1985)

Id.: B3 ; α, δ (1950) ; S(408MHz)

Catalogue  References  Contents  Remarks

The Parkes-MIT-NRAO (PMN) Surveys 4.5Mb 50814 sources

VIII/38/

N-T-f-v Griffith M.R., Wright A.E., Burke B.F., Ekers R.D., Astrophys. J. Suppl. Ser. 90, 179 (1994) Wright A.E., Griffith M.R., Burke B.F., Ekers R.D., Astrophys. J. Suppl. Ser. 91, 111 (1994) Griffith M.R., Wright A.E., Burke B.F., Ekers R.D., Astrophys. J. Suppl. Ser. 97, 347 (1995) Wright A.E., Griffith M.R., Hunt A.J., Troup E., Burke B.F., Ekers R.D., Astrophys. J. Suppl. Ser. 103, 145 (1996)

Id.: PMN (J) α, δ (J2000) S(4850MHz) + σ Notes and Remarks Supersedes VIII/27 and J/ApJS/97/347 δ < 10^{∘}

Composite CO Survey of the Milky Way 71Mb 720 vel. maps

VIII/39/

M-N-T-f Dame T.M., Ungerechts H., Cohen R.S., de Geus E., Grenier I.A., May J., Murphy D.C., Nyman L.-A., Thaddeus P., Astrophys. J. 322, 706 (1987)

720 latitude-velocity maps (FITS files), one for each 30’ of galactic longitude One Velocity-integrated map by exchange from NASA-Goddard

The GB6 catalog of radio sources 22Mb 75162 sources

VIII/40/

T-f-v Gregory P.C., Scott W.K., Douglas K., Condon J.J., Astrophys. J. Suppl. Ser. 103, 427 (1996) α, δ (B1950, J2000) ℓ, b F (4.85GHz) + σ FWHM axises Major axis PA Local sky level Map pixel position

Supersedes VIII/14/ δ > 0^{∘}

The Texas survey of radio sources at 365MHz 10Mb 66841 sources

VIII/42/

T-f-v

Douglas  J.N.,  Bash

F.N., Bozyan F.A., Torrence G.W., Wolfe C., Astron. J. 111, 1945 (1996)

Id.: TXS

α, δ (1950) +σ

365MHz flux, σ, S/N

Spectral index Structure Overlapping comments Simpler model results for xxx sources Zone −35.5^{∘} < δ < 71.5^{∘} from AAS CD-ROM#6

A radio survey of clusters of galaxies 197Kb 150 clusters

VIII/43/

T-f-v Andernach H. et al., Haslam C.G.T. et al, Waldthausen H. et al, Reuter H.-P. & Andernach, Astron. Astrophys. Suppl. Ser. 31, 99; 36, 237; 41, 339; 43, 155; 65, 561 ; 73, 265 ; 82, 279 Id.: ACO α, δ (1950) Optical data Fluxes around 2.7, 4.8, 10.7 GHz Spectral indices References

The Miyun 232MHz survey 2.7Mb 34426 sources

VIII/44/

T-f-v Zhang X., Zheng Y., Chen H., Wang S., Cao A., Peng B., Nan R., Astron. Astrophys. Suppl. Ser. 121, 59 (1997)

Id.: Miyun α, δ (1950) S_{peak}, S_{int}, S/N

Catalogue  References  Contents  Remarks

Radio positions and Optical Identifications for a sample of Southern flat-spectrum radio sources 59Kb 198 sources

VIII/45/

T-f-v WHITE G.L., BATTY M.J., BUNTON J.D., BROWN D.R., CORBEN J.B., Mon. Not. R. Astron. Soc. 227, 705 ; 248, 398 Id.: PKS-system α, δ (1950): radio & optical S(2.7GHz), spectral index Optical counterpart source type

Galactic 21-cm line Survey 147Mb 58 FITS cubes

VIII/47/

N-T-f Westerhout G., Wendlandt H.U., Astron. Astrophys. Suppl. Ser. 49, 137 (1982)

58 FITS files with flux densities by exchange from NASA-Goddard

The Zelenchuk survey 667Kb 12551 sources

VIII/49/

T-f-v Amirkhanyan V.R., et al., The Zelenchuk 3.9 GHz Survey (1989) Amirkhanyan V.R., et al., Pis’ma Astron. Zh.18, 396 (1992) Amirkhanyan V.R., et al. , Astron. Zh.69, 225 (1992) Larianov M.G., et al., Astron. Astrophys. Suppl. Ser. 106, 119 (1994) Gorshkov A.G., Konnikova V.K., Astron. Zh.72, 291 (1995) (Astron. Rep. 39, 257) Larianov M.G., Soobshch. Spets. Astrof. Obs., 68, 14 (1991)

Id.: Zel, other α, δ (1950, 2000) Flux densities at 3.9GHz

{Zones  −1  to \atop 14^{∘}}

VLA Survey of Rich Clusters of Galaxies 149Kb 1037 sources

VIII/50/

T-f-v Slee O.B., Roy A., Savage A., Aust. J. Phys. 47, 145 (1994) Slee O.B., Roy A., Andernach H. , Aust. J. Phys. 49, 977 (1996)

Id.: Cluster designation α, δ (1950) F (1.5 and 4.9GHz) Size of source z (of cluster)

from H. Andernach

The MIT-Green Bank (MG) 5 GHz Survey 1.6Mb 20753 sources

VIII/52B/

T-f-v Bennett C.L., Lawrence C.R., Burke B.F., Hewitt J.N., Mahoney J., Astrophys. J. Suppl. Ser. 61, 1 (1986) Langston G.I., Heflin M.B., Conner S.R., Lehar J., Carrilli C.L., Burke B.F., Astrophys. J. Suppl. Ser. 72, 621 (1990) Griffith M., Langston G., Heflin M., Conner S., Lehar J., Burke B., Astrophys. J. Suppl. Ser. 74, 129 (1990)

Id.: MG α, δ (1950) S and S/N at 4830 or 4775 MHz Spectral index

from ADS

The Green Bank Survey 203Kb 3784 sources

VIII/53/

T-f-v Maslowski J. , Acta Astron. 22, 227 (1972) Machalski J. , Acta Astron. 28, 367 (1978) Rys S., Machalski J. , Acta Astron. 37, 163 (1978)

Id.: GB α, δ (1950) S(1400MHz) Extension flags

from ADS

Atlas of Galactic Neutral Hydrogen 235Mb 721 maps

VIII/54/

T-f-v Hartmann D., Burton W.B., Cambridge University Press (1997)

21-cm HI maps in (ℓ, b), step 0.5^{∘} by exchange from NASA-Goddard

142  Catalogues available at CDS  ( January 19, 2008)

Catalogue  References  Contents  Remarks

DRAO Penticton P-survey Database (May 1990) 165Kb 1692 sources

VIII/55/

T-f-v — , Compilation of 22 lists Id.: IAU α, δ (1950) +σ ν, F (ν), FWHM Obs. date Comments

From ADS

Compendium of Radio Measurements of Bright Galaxies 926Kb 7156 sources

VIII/56/

T-f-v Haynes R.F., Huchtmeier W.K.H., Siegman B., Wright A.E., CSIRO Publication (1975)

Id.: IAU

α, δ (1950) +σ

ν, F (ν)

Angular size UBV, V_{r} References From ADS Catalogues

Fluxes of Faint Radio Sources at 2.7/4.75 GHz (32.8d − −33.5^{∘}) 28Kb 239 sources

VIII/57/

T-f-v Forkert T., Altschuler D.R., Astron. Astrophys. Suppl. Ser. 70, 77 (1987)

Id.: IAU, other α, δ (1950) Fluxes at 2695 and 4750 MHz, mean errors Spectral index

Revised Version

A Survey of Radio H II Regions in the Northern Sky 44Kb 462 sources

VIII/58/

N-T-f-v Lockman F.J., Astrophys. J. Suppl. Ser. 71, 469 (1989)

Id.: ; α, δ (1950); ℓ, b ; T, σ(T), FWHM ; V_{LSR}

Interferometer Phase Calibration Sources 175Kb 2921 sources

VIII/60/

T-f-v Patnaik A.R., Browne I.W.A., Wilkinson P.N., Wrobel J.M., Mon. Not. R. Astron. Soc. 254, 655 (1992) ; Mon. Not. R. Astron. Soc. 293, 257; Mon. Not. R. Astron. Soc. 300, 790 Id.: Bname, Jname α, δ (J2000) F (8.4GHz) Comments

The NAIC 611 MHz Multibeam Sky Survey Source List 248Kb 3122 sources

VIII/61/

T-f-v Durdin J.M., Pleticha D., Condon J.J., Yerbury M.J., Jauncey D.L., Hazard C., National Astronomy and Ionosphere Center Report 45 (1975)

Id.: IAU Name α, δ (B1950) F (611MHz) + σ Source size by exchange from NASA-Goddard Zone 0 < α < 12h, −2^{∘} < δ < 18^{∘}

The Westerbork Northern Sky Survey 25Mb 229420 sources

VIII/62/

f-v Netherlands Foundation for Research in Astronomy ((NFRA/ASTRON), Leiden Observatory (1998)

Id.: IAU

α, δ (B1950, J2000) Source type & extension S_{p}(330MHz), S_{i}(330MHz) rms noise

Catalogue  References  Contents  Remarks

The ATESP Survey 295Kb 3370 sources

VIII/63/

T-f-v Prandoni I., Gregorini L., Parma P., de Ruiter H.R., Vettolani G., Wieringa M.H., Ekers R.D., , Astron. Astrophys. Suppl. Ser. 146, 41 (2000)

Id.: ATESP α, δ (J2000) S(1.4GHz), S_{p}(1.4GHz) FWHM, flags

Westerhout’s Catalogue of 82 Discrete Sources 13Kb 107 sources

VIII/64/

T-f-v Westerhout G., Bull. Astron. Inst. Netherlands 14, 215 (1958)

Id.: W α, δ (J2000) Notes

The NRAO VLA Sky Survey 271Mb 1773484 sources

VIII/65/

f-v Condon J.J., Cotton W.D., Greisen E.W., Yin Q.F., Perley R.A., Taylor G.B., Broderick J.J., Astron. J. 115, 1693 (1998)

Id.: NVSS α, δ (J2000) Dimensions S(1.4GHz), residual flux Polarized flux

IRAM keyproject “Small-scale structure of pre-star forming regions” 69Mb 40 data-cubes

VIII/66/

T-f-v Falgarone E., Panis J.-F., Heithausen A., Perault M., Stutzki J., Puget J.-L., Bensch F. , Astron. Astrophys. 331, 669 (1998) Bensch F., Panis J.-F., Stutzki J., Heithausen A., Falgarone E. , Astron. Astrophys. 365, 275 (2001)

Id.: FITSfile

ℓ, b of obs. window Velocity range FITS data cubes (αδV)

Catalogue of scintillating radio sources 31Kb 414 records

VIII/67/

f-v Artyukh V.S., Tyul’bashev S.A., Isaev E.A., Astron. Zh. 73, 661 (1996); Astron. Zh. 75, 323 (1998)

Id.: 7C α, δ (B1950) S(102MHz) Notes From Moscow Data Center

The WISH catalogue at 352 MHz 10Mb 90357 sources

VIII/69A/

f-v Carlos De Breuck, Yuan Tang, Ger de Bruyn, Huub Rottgering, Wil van Breugel, Roeland Rengelink, Astron. Astrophys. 394, 59 (2002)

Id.: WISH

α, δ (B1950, J2000) Source dimensions S (peak, integrated) Flags See also VIII/62 (WENSS)

The FIRST Survey Catalog, Version 2003 April 99Mb 811117 sources

VIII/71/

f-v Becker R.H., Helfand D.J., White R.L., Gregg M.D., Laurent-Muehleisen S.A., Astrophys. J. 475, 479 (1997)

Id.: FIRST, Field Name α, δ (J2000) F_{peak}, F_{int}, S/N Source size Warning flag Supersedes VIII/59

Catalogue  References  Contents  Remarks

CLASS (Cosmic Lens All-Sky Survey) of radio sources 2.9Mb 23418 sources

VIII/72/

f-v Myers S.T., Jackson N.J., Browne I.W.A., de Bruyn A.G., Pearson T.J., Readhead A.C.S., Wilkinson P.N., Biggs A.D., Blandford R.D., Fassnacht C.D., Koopmans L.V.E., Marlow D.R., McKean J.P., Norbury M.A., Phillips P.M., Rusin D., Shepherd M.C., Sykes C.M. , Mon. Not. R. Astron. Soc. 341, 1 (2003)

Id.: GB6

α, δ (J2000) (pointing, components) S(8.46GHz) a, b/a, PA

0^{∘} < δ < 75^{∘}, |b| ≥ ; 10^{∘} ; α ≥ −0.5

The HI Parkes All Sky Survey Catalogue 2.0Mb 4315 sources

VIII/73/

f-v Meyer, M.J., Zwaan, M.A., Webster, R.L., et al. , Mon. Not. R. Astron. Soc. 350, 1195 (2004)

Id.: HIPASS α, δ (J2000) V at centre of profile, clipped W (profile width) S_{p}m S_{int}

Master Catalog of radio sources 122Mb 375105 radio-sources

VIII/74A/

f-v B. Vollmer , Astron. Astrophys. 431, 1177 (2005)

Id.: NVSS, TXS, ...

α, δ (J2000)

a, b in log F = a + b log ν Individual fluxes

A high sensitivity HI survey of the sky at δ < +25^{∘} 241Mb 50980 profiles

VIII/75/

f-v Bajaja E., Arnal E.M., Larrarte J.J., Morras R., Poeppel W.G.L., Kalberta P.M.W., Astron. Astrophys. 440, 676 (2005)

3-D cube of H I data

50980 profiles on a galactic grid Δℓ = 0.5^{∘}, Δb = 0.5^{∘}/ cos b

Argentine survey

Leiden/Argentine/Bonn (LAB) Survey of Galactic H I 2136Mb 721×360 spectra

VIII/76/

f-v Kalberla P.M.W., Burton W.B., Hartmann D., Arnal E.M., Bajaja E., Morras R., Poeppel W.G.L. , Astron. Astrophys. 440, 775 (2005) ℓ, b (J2000) N_{H}(ℓ, b) for velocities in ranges ±25, ±50, ±75, ±100, ±125, ±200 and ±400 km/s 3-D maps as FITS files (B, b) maps as FITS files

Supersedes VIII/54

A digital archive of H I 21cm line spectra of optically targeted galaxies 1.4Mb 8850 galaxies

VIII/77/

f-v Springob C.M., Haynes M.P., Giovanelli R., Kent B.R., Astrophys. J. Suppl. Ser. 160, 149 (2005)

Id.: UGC/AGC, Other

α, δ (J2000) (a, b), T V_{r}, HI flux, width Telescope Notes

The VLA Low-frequency Sky Survey 7.1Mb 68308 sources

VIII/79/

f-v Cohen A.S., Lane W.M., Cotton W.D., Kassim N.E., Lazio T.J.W., Perley R.A., Condon J.J., Erickson W.C., Astron. J. 134, 1245 (2007)

Id.: VLSS α, δ (J2000) S_{p}(74MHz) Source dimensions

{74MHz (4m) \atop δ ≥ −30^{∘}}

Catalogue  References  Contents  Remarks

The UTR-2 Very Low-Frequency Sky Survey Data 1.1Mb 2302 sources

VIII/80/

f-v Braude et al., Astrophys. Space Sci. 54, 37 (1978); 64, 73 (1979) ; 74, 409 (1981) ; 111, 1 (1985); 213, 1 (1994); 280, 235 (2002)

Id.:

α, δ (J2000)

S(16.7MHz) with σ Spectral Index Quality Index S at several frequencies between 10 and 20 MHz From Kharkov Institute of Rsdio Astronomy Supersedes VIII/41

The Sydney University Molonglo Sky Survey (SUMSS, Version 2.0) 27Mb 210412 sources

VIII/81/

f-v Mauch T., Murphy T., Buttery H.J., Curran J., Hunstead R.W., Piestrzynski B., Ropbertson J.G., Sadler E.M., Mon. Not. R. Astron. Soc. 342, 1117 (2003)

Id.: SUMSS α, δ (J2000) + σ S_{peak}, S_{tot} + σ at 843MHz

Fitted/deconvolved major/minor axes Supersedes VIII/78 δ < −30^{∘}, |b| > 10^{∘}

The Molonglo Galactic Plane Survey (MGPS-2): Compact Source Catalogue 8.6Mb 48850 sources

VIII/82/

f-v Murphy T., Mauch T., Green A., Hunstead R.W., Piestrzynska B., Kels A.P., Sztajer P. , Mon. Not. R. Astron. Soc. 382, 382 (2007)

Id.: MGPS

α, δ (J2000) + σ

S_{peak}, S_{tot} + σ at 843MHz

Fitted/deconvolved

major/minor

axises

IX. High Energy Data

Catalogue  References  Contents  Remarks

List of Positions of all X-ray sources with positions known more accurately than those given in the 4U or 2A catalogues (2nd Ed.) 126Kb 1257 sources

IX/6/

B-I-M-N-T-f Dolan J.F., NASA/GSFC (1983)

Id.: 4U, X-ray, optical

α, δ (B1950) X-ray flux Notes and remarks Was numbered VII/55

Second Catalogue of X-ray Sources 216Kb 677 sources

IX/8/

B-M-N-T-f-v Amnuel P.R., Gusenov O.H., Rakhaminov Sh.Yu., Astrophys. Sp. Sci. 82, 3 (1982)

Id.: Names α, δ (1950); ℓ, b Area of the error box Dates, Measuring system Energy range I(max), I(min) Comments Keypunched at the CDS Was previously VII/81

Catalog of SAS-2 Gamma-Ray Observations 26Kb 183 sources

IX/9/

N-T-f Fichtel C.E., Hartman R.C., Hunter S.D., Kniffen D.A., Thompson D.J., Ögelman H.B., Tümer T., Özel M.E., Laboratory for High Energy Astrophysics, NASA Goddard Space Flight Center (1990)

Id.: Galaxy name, PSR Object type α, δ (1950) Energy Notes and References

{Was \atop VII/130/}

The ROSAT All-Sky Bright Source Catalogue 95Mb 18811 sources

IX/10A/

M-N-T-f-v Voges W., Aschenbach B., Boller Th., Braeuninger H., Briel U., Burkert W., Dennerl K., Englhauser J., Gruber R., Haberl F., Hartner G., Hasinger G., Kuerster M., Pfeffermann E., Pietsch W., Predehl P., Rosso C., Schmitt J.H.M.M., Truemper J., Zimmermann H.-U. , Astron. Astrophys. 349, 389 (1999)

Id.: 1RXS α, δ (2000), σ Count rate +σ Exposure time Hardness ratios Source extent Flags Correlation with other catalogues from MPE Garching See also IX/29

EINSTEIN Observatory catalog of IPC X-ray sources 675Kb 4809 sources

IX/13/

T-f-v Harris D.E., Forman W, Gioa I.M., Hale J.A., Harnden Jr. F.R., Jones C., Karakashian T., Maccacaro T., McSweeney J.D., Primnini F.A., Schwarz J., Tananbaum H.D., Thurman J. , SAO HEAD CD-ROM Series I (Einstein), Nos 18-36 (1994)

Id.: ID, 2E (from position) α, δ (1950), σ Count rate + σ Background counts S/N Remarks & References

The Einstein Observatory Extended Medium-Sensitivity Survey (EMSS) 170Kb 835 sources

IX/15/

N-T-f-v Gioia I.M., MacCacaro T., Schild R.E., Wolter A.: 1990 , Astrophys. J. Suppl. Ser. 72, 567 Stocke J.T., et al.: 1991 , Astrophys. J. Suppl. Ser. 76, 813 Id.: MS, optical/radio α, δ (1950), σ Optical position offset F_{X} with S/N F (radio) Count rate, net counts Exposure time m_{V} , log(F_{X}/F_{V} ) z, class, comments Previously VII/152

Catalogue  References  Contents  Remarks

EINSTEIN extended source survey (EXSS) 434Kb 1326 sources

IX/18/

T-f-v Oppenheimer B.R., Helfand D.J., Gaidos E.J., Astron. J. 113, 2134 (1997)

Id.: EXSS α, δ (1950) IPC Count rates + S/N Aperture offsets Coincence with PPM, EMSS, ZwCl, Abell lists

X-ray lightcurves of bright AGNs from Ginga 1.1Mb 8 lightcurves

IX/19/

N-T-f Hayashida K., Miyamoto S., Kitamoto S., Negoro H., Inoue H., Astrophys. J. 500, 642 (1998)

X-ray light curves for NGC 4051, 4151, 5506, 5548; 3C 273; MCG-6-30-15; ESO 103-G035; PKS 2155-304 by exchange from NASA-Goddard

The Fourth Revised BATSE Burst Catalog 1.2Mb 2702 bursts

IX/20A/

T-f-v Paciesas W.S., Meegan C.A., Pendleton G.N., Briggs M.S., Kouveliotou C., Koshut T.M., Lestrade J.P., McCollough M.L., Brainerd J.J., Hakkila J., Henze W., Preece R.D., Connaughton V., Kippen R.M., Mallozzi R.S., Fishman G.J., Richardson G.A., Sahi M. , Astrophys. J. Suppl. Ser. 122, 465 (1999)

Id.: 4B trigger number Trigger time α, δ (J2000) ; ℓ, b Error box Overwrite flags Fluence data Comments Ulysses boxes

The Einstein Extended Deep Survey 874Kb 388 sources

IX/22/

T-f-v Primini F.A., Murray S.S., Huchra J., Schild R., Burg R., Giacconi R., Astrophys. J. 374, 440 (1991)

Id.: HRI, IPC α, δ (1950) Count rates Optical Counterparts Cross-Reference HRI/IPC Target Information

from ADS

The HRI Source List 249Kb 600 sources

IX/24/

f-v

Id.: Sequence α, δ (1950), σ Count rate, S/N from ADS

Einstein Obs. IPC Parameters 2.9Mb 4295 obs.

IX/25/

f-v Giacconi et al., ???

Id.: IPC α, δ (1950) Start/Stop Times

{From \atop ADS}

ROSAT Results Archive Sources for the HRI 43Mb 77424 sources

IX/28A/

T-f-v ROSAT Scientific Team, MPE, Garching bei München Id.: 1RXH α, δ (2000) ; ℓ, b Count Rate, σ, S/N t_{exp}, Obs. date Processing details, notes

Catalogue  References  Contents  Remarks

ROSAT All-Sky Survey Faint source Catalogue (RASS-FSC)

22Mb 105924 sources

IX/29/

T-f-v Voges W., Aschenbach B., Boller Th., Brauninger H., Briel U., Burkert W., Dennerl K., Englhauser J., Gruber R., Haberl F., Hartner G., Hasinger G., Pfeffermann E., Pietsch W., Predehl P., Schmitt J., Trumper J., Zimmermann U. Max-Planck-Institut fur extraterrestrische Physik, , IAU Circ. 7432 (2000) ; Max-Planck-Institut fur extraterrestrische Physik, Garching (2000)

Id.: 1RXS α, δ (J2000) Count rate +σ Exposure time Hardness ratios Source extent Flags from MPE Garching See also IX/10

The Second ROSAT Source Catalog of Pointed Observations 54Mb 95331 sources

IX/30/

T-f-v ROSAT Consortium, ROSAT News 72, 25-May-2000 Id.: ROSAT Seq α, δ (2000) ; ℓ, b Positional accuracies Count Rates Hardness ratios Exposure time Extent size

Supersedes IX/11

The WGACAT version of the ROSAT PSPC Catalogue, Rev. 1 85Mb 68907 sources

IX/31/

T-f-v White N.E., Giommi P., Angelini L., HEASARC, Code 660, LHEA/GSFC, Greenbelt Id.: 1WGA, 1RXP α, δ (2000) ; ℓ, b (α, δ) of image, target Positional accuracies Count Rates Softness & Hardness ratios Exposure time S/N, variability χ^{2} Counterparts in X-ray, Optical, IR, Radio See also IX/11, IX/29, IX/30 Supersedes IX/12

The ROSAT Bright Survey 335Kb 2072 sources

IX/32/

f-v Fischer J.-U., Hasinger G., Schwope A.D., Brunner H., Boller T., Trumper J., Voges W., Neizvestny S., Astron. Nachr. 319, 347 (1998); 321, 1 (2000)

Id.: HS, RBS, other

α, δ (J2000) F_{X}, F_{R} z, Classification Optical Counterparts Sources brighter than 0.2ct/s

The Brera Multi-scale Wavelet ROSAT HRI source catalogue

37Mb 29089 sources

IX/34/

f Panzera M.R., Campana S., Covino S., Lazzati D., Mignani R.P., Moretti A., Tagliaferri G., Astron. Astrophys. 399, 351 (2003)

Id.: 1BMW, FIRST, GSC2, 2MASS, IRAS α, δ (J2000) \underline{X-ray data}: HRI parameters, counts, fluxes in ROSAT bands, extension \underline{FIRST data}: extensions, 21-cm flux \underline{GSC2S data}: extensions, photographic JV F N, μ \underline{2MASS data}: JHK photometry \underline{IRAS data}: IRAS fluxes

Second EUVE Catalogues 284Kb 1196 sources

IX/35/

f-v Bowyer S., Lampton M., Lewis J., Wu X., Jelinsky P., Malina R.F. , Astrophys. J. Suppl. Ser. 102, 129 (1996) Lampton M., Lieu R., Schmitt J.H.M.M., Bowyer S., Voges W., Lewis J., Wu X., Astrophys. J. Suppl. Ser. 108, 545 (1997) Christian D.J., Craig N., Cahill W., Roberts B., Malina R.F. , Astron. J. 117, 2466 (1999)

Id.: EUVE α, δ (2000) Count rates at 10, 20, 40 and 60nm Optical counterparts

Supersedes II/203

Granat/Phebus catalog of cosmic gamma-ray bursts 32Kb 210 sources

IX/36/

f-v Terekhov O.V., Tkachenko A.Yu, Denisenko D.V., Lobachev V.A., Sunyaev R.A., Kovtun,A.V., Kuznetsov A.V., Barat C., Dezalay J.-P., Talon R., Vedrenne G. , Pis’ma Astron. Zh.20, 323 (1994); 21, 83 (1995), 24, 833 (1998); 28, 403 (2002)

Id.: Burst name

Date and time of Burst Fluence, peak flux Spectral hardness Fitting parameters Notes

150  Catalogues available at CDS  ( January 19, 2008)

Catalogue  References  Contents  Remarks

The XMM-Newton Serendipitous Source Catalogue 333Mb 56711 sources

IX/37/

f-v XMM-Newton Survey Science Centre Consortium , XMM-SSC, Leicester, UK Id.: 1XMM α, δ (J2000) Observation Date X-ray counts & fluxes in enery ranges: 0.2–0.5keV, 0.5–2.0keV, 2.0–4.5keV, 4.5–7.5keV, 7.5–12.0keV Cross-correlation products

The XMM-Newton 2nd Serendipitous Source Catalogue 466Mb records

IX/39/

f-v XMM-Newton Survey Science Centre Consortium , XMM-SSC, Leicester, UK Id.: 2XMM α, δ (J2000) Observation Date X-ray counts & fluxes in energy ranges: 0.2–0.5keV, 0.5–2.0keV, 2.0–4.5keV, 4.5–7.5keV, 7.5–12.0keV Supersedes IX/37

J. Published Data by Journal

Tables available from the VizieR service are marked V

J/A+A/8/315 See II/260/ J/A+A/17/47 See II/63A/ J/A+A/18/169 See VII/15/ J/A+A/18/341/ catalog.dat notes.dat (71Kb)

Stock J., Wroblewski H.

Astron. Astrophys., 18, 341-349 (1972) Double stars of equal components.

J/A+A/18/428 See III/22/ J/A+A/35/393/ table4.dat (15Kb)

Katgert J.K.

Astron. Astrophys. 35, 393 (1974)

A Second Deep Survey with the Westerbork telescope at 1415 and 610 MHz: an attempt to obtain deep identifications.

J/A+A/39/159 See II/60/ J/A+A/39/461 See III/247/ J/A+A/46/333 See VIII/7A/ J/A+A/60/221 See VII/50/ J/A+A/78/687 See III/46/ J/A+A/82/253 See II/63A/ J/A+A/85/329/ table2 (12Kb)

Pauliny-Toth I.I.K., Steppe H., Witzel A.

Astron. Astrophys. 85, 329 (1980)

A deep survey of selected regions for extragalactic sources at 4.85GHz J/A+A/108/L7 See VII/56/ J/A+A/108/256 See VII/32/ J/A+A/109/181 See VI/87/ J/A+A/114/278 See VI/81/ J/A+A/123/39 See V/23/ J/A+A/124/50 See VI/79/ J/A+A/125/359 See III/231/ J/A+A/130/191 See I/169/ J/A+A/171/261/ table1.dat (57Kb)

Caswell J.L., Haynes R.F.

Astron. Astrophys. 171, 261 (1987)

Southern HII Regions: an extensive study of radio recombination line emission J/A+A/186/1 See VI/47/ J/A+A/190/342 See VI/79/ J/A+A/202/309 See VI/81/ J/A+A/214/68 See II/144/ J/A+A/222/69 See II/180/ J/A+A/227/82/ table1a.dat table3.dat (46Kb)

Epchtein N., Le Bertre T., Lepine J.R.D.

Astron. Astrophys. 227, 82 (1990)

Carbon star envelopes: near-IR photometry, mass loss and evolutionary status of a sample of IRAS stars. J/A+A/230/307 See III/212/

J/A+A/235/185/  table2  (9Kb)

Motch C., Hudec R., Christian C. Astron. Astrophys. 235, 185 (1990) Quiescent candidates for the repeating optical transient OT 1809+31

J/A+A/240/262/  table3.dat  (50Kb)

Aparicio A., Bertelli G., Chiosi C., Garcia-Pelayo J.M. Astron. Astrophys. 240, 262 (1990) CCD UBVR photometry of the old rich open cluster King 2: comparison with theoretical models.

J/A+A/241/551/  table2.dat table3.dat table4.dat table5.dat  (28Kb)

Wendker H.J., Higgs L.A., Landecker T.L. Astron. Astrophys. 241, 551 (1991) The Cygnus X region. XVIII. A detailed investigation of radio continuum structure on large and small scales

J/A+A/242/401/  wx-hyi.dat  (306Kb)

Kuulkers E., Hollander A., Oosterbroeck T., van Paradijs J. Astron. Astrophys. 242, 401 (1991) Walraven photometry of WX Hyi during quiescence and outburst.

J/A+A/243/386/  table1 table2  (45Kb)

Schwan H. Astron. Astrophys. 243, 386 (1991) The distance and main sequence of the Hyades cluster based on 145 stars with highly accurate proper motions obtained from work on the catalogues FK5 and PPM.

J/A+A/250/370/  appena appenc append appene appenf table3  (25Kb)

Bixler J.V., Bowyer S., Laget M. Astron. Astrophys. 250, 370 (1991) A high galactic latitude survey of far-ultraviolet excess objects

J/A+A/256/121 See III/191/

J/A+A/258/217 See I/197A/

J/A+A/258/243/  table1.dat  (9Kb)

Proust D., Quintana H., Mazure A., da Souza R., Sodre L., Capelato H.V. Astron. Astrophys. 258, 243 (1992) The cluster of galaxies Abell 151.

J/A+A/258/423/  table1.dat  (36Kb)

Muratorio G., Viotti R., Friedjung M., Baratta G.B., Rossi C. Astron. Astrophys. 258, 423-431 (1992) The spectrum of the VV Cep star KQ Puppis (Boss 1985). IV. The Fe II spectrum, level population and spectral synthesis

J/A+A/264/557 See VI/77/

J/A+A/265/32/  fields  (59Kb)

Robin A.C., Creze M., Mohan V. Astron. Astrophys. 265, 32 (1992) The radial structure of the galactic disc

J/A+A/265/77/  tt-ari.dat  (65Kb)

Hollander A., van Paradijs J. Astron. Astrophys. 265, 77 (1992) Quasi-periodic oscillations in TT Ari

J/A+A/267/515/  table1 table2 table3 table4 table5 tablea1 tablea2  (57Kb)

Omont A., Loup C., Forveille T., te Lintel Hekkert P., Habing H., Sivagnanam P. Astron. Astrophys. 267, 515, (1993) Characterization and proportion of very cold C-rich circumstellar envelopes

J/A+A/268/419/  table3  (15Kb)

Courtes G., Petit H., Hua C.T., Martin P., Blecha A., Huguenin D., Golay M. Astron. Astrophys. 268, 419 (1993) Structure of the spiral arms of NGC 4258 in Halpha and at 2000Angstroems.

J/A+A/270/522 See VI/63/

J/A+A/271/515/  beta.dat gamma.dat  (15Kb)

Smalley B., Dworetsky M.M. Astron. Astrophys. 271, 515 (1993)

The atmospheric parameters of A and F stars. I. Comparison of various methods

J/A+A/272/235/  table1.dat table2.dat table3.dat  (11Kb)

Kun M., Prusti T. Astron. Astrophys. 272, 235 (1993) Star formation in L1251: distance and members

J/A+A/274/105/  table1 table2 tables  (29Kb)

WALTER R., FINK H. H. Astron. Astrophys. 274, 105 (1993) The ultraviolet to soft X-ray bump of Seyfert 1 type Active Galactic Nuclei

J/A+A/274/335/  table1 table2  (14Kb)

Smith K.C., Dworetsky M.M. Astron. Astrophys. 274,335 (1993) Elemental abundances in normal late-B and HgMn stars from co-added IUE spectra. I . Iron-peak elements

J/A+A/274/555/  table1.dat  (13Kb)

Stathopoulou M., Alissandrakis C.E. Astron. Astrophys. 274, 555 (1993) A study of the asymmetry of Fe I lines in the solar spectrum

J/A+A/274/895/  table2.dat table3.dat  (17Kb)

Van de Steene G.C.M., Pottasch S.R. Astron. Astrophys. 274, 895 (1993) Radio continuum observations of southern planetary nebulae candidates

J/A+A/275/67/  e1  (19Kb)

Brand J., Blitz L. Astron. Astrophys. 275, 67 (1993) The Velocity Field of the outer Galaxy

J/A+A/275/101/  stars table8  (46Kb)

Edvardsson B., Andersen J., Gustafsson B., Lambert D.L., Nissen P.E., Tomkin J. Astron. Astrophys. 275, 101 (1993) The chemical evolution of the galactic disk I. Analysis and results

J/A+A/275/163/  table1.dat table2.dat table3.dat table4.dat table5.dat  (26Kb)

Kastner J.H., Forveille T., Zuckerman B., Omont A. Astron. Astrophys. 275, 163 (1993) Probing the AGB tip: luminous carbon stars in the galactic plane

J/A+A/275/201/  table1.dat table2.dat  (10Kb)

Echevarria J., Costero R., Michel R. Astron. Astrophys. 275, 201 (1993) Stroemgren photometry of dwarf novae

J/A+A/275/484/  table1 table2 table3a table3b  (10Kb)

Zsoldos E., Fernie J.D., Ferro A. Arellano, Seager S. Astron. Astrophys. 275, 484 (1993) The double-mode semiregular variable UU Herculis: 1990-1992 photometry

J/A+A/275/549/  table  (14Kb)

Jenniskens P., Greenberg J.M. Astron. Astrophys. 275, 549 (1993) Environment dependence of interstellar extinction curves

J/A+A/276/25/  table1 table2 table3 table4  (21Kb)

Israel F.P., Johansson L.E.B., Lequeux J., Booth R.S., Nyman L.-A., Crane P., Rubio M., de Graauw T., Kutner M.L., Gredel R., Boulanger F., Garay G., Westerlund B.E. Astron. Astrophys. 274, 25 (1993) Results of the ESO-SEST Key Programme on CO in the Magellanic Clouds. I. A survey of CO in the LMC and the SMC.

J/A+A/276/142/  table2  (16Kb)

Hill G.M., Landstreet J.D. Astron. Astrophys. 276, 142 (1993) Compositional differences among the A-type stars. I. Six narrow lined stars.

J/A+A/276/345/  table1 table3 table4  (25Kb)

Jestu L. Astron. Astrophys. 276, 345 (1993)

A decade of photometry of LQ Hydrae

J/A+A/276/389/  table2  (52Kb)

Kipper T., Kipper M. Astron. Astrophys. 276, 389 (1993) The spectrum of FG Sge in 1992

J/A+A/277/453/  refs.dat table1.dat table2.dat  (24Kb)

David P., Le Squeren A.M., Sivagnanam P. Astron. Astrophys. 277, 453 (1993) An OH satellite line maser survey of cool IRAS sources and circumstellar envelope evolution

J/A+A/278/6/  table1 table2 table3 table4 table5 table6  (26Kb)

Xie G. Z., Zhang Y.H., FaN J.H., LIU F.K. Astron. Astrophys. 278, 6 (1993) The relation between BL Lac objects and OVV quasars, and the unified model of BL Lac objects, FR-I and FR-II(G) radio galaxies

J/A+A/278/36/  table1  (9Kb)

Magnier E.A., Battinelli P., Lewin W.H.G., Haiman Z., Van Paradijs J., Hasinger G., Pietsch W., Supper R., Trumper J. Astron. Astrophys. 278, 36 (1993) Automated identification of OB associations in M 31

J/A+A/278/379/  table1  (20Kb)

Briel U.G., Henry J.P. Astron. Astrophys. 278, 379 (1993) X-ray Emission from a Complete Sample of Abell Clusters of Galaxies

J/A+A/278/449/  table1 table2 table3 table4  (26Kb)

JETSU L., PELT J., TUOMINEN I. Astron. Astrophys. 278, 449 (1993) Spot and flare activity of FK Comae Berenices: long-term photometry

J/A+A/279/61/  tables  (17Kb)

NORGAARD-NIELSEN H.U., GOUDFROOIJ P., JORGENSEN H.E., HANSEN L. Astron. Astrophys. 279, 61 (1993) The extinction and star clusters in NGC 1275

J/A+A/279/541/  table1 table2 table3  (14Kb)

Van Dishoeck E.F., Jansen D.J., Phillips T.G. Astron. Astrophys. 279, 541 (1993) Submillimeter observations of the shocked molecular gas associated with the supernova remnant IC 443

J/A+A/279/567/  table2.dat table3.dat table4.dat  (15Kb)

de Freitas Pacheco J.A., Costa R.D.D., Maciel W.J. Astron. Astrophys. 279, 567 (1993) Abundances of non-type I planetary nebulae in the LMC

J/A+A/280/365/  table1  (20Kb)

LE COARER E., ROSADO M., GEORGELIN Y., VIALE A., GOLDES G. Astron. Astrophys. 280, 365 (1993) H-alpha Survey of the Small Magellanic Cloud

J/A+A/280/581/  refs.dat table1.dat table2.dat table3.dat table4.dat  (30Kb)

Pasquali A., Perinotto M. Astron. Astrophys. 280, 581 (1993) Chemical behaviour of planetary nebulae and galactic abundance gradients

J/A+A/281/L25 See I/207/

J/A+A/281/15/  table1 table2  (14Kb)

RYS S. Astron. Astrophys. 281, 15 (1994) The asymmetries in radio source structures. I. A comparison of two classical models.

J/A+A/281/161/  table4 table5 table6 table7 table8 table9 table10  (26Kb)

ALTENHOFF W.J., THUM C., WENDKER H.J. Astron. Astrophys. 281, 161 (1994) Radio emission from stars: a survey at 250 GHz

J/A+A/281/355/  table2  (43Kb)

BRINKMANN W., SIEBERT J., BOLLER T.

Astron. Astrophys. 281, 355 (1994) The X-ray AGN content of the Molonglo 408 MHz survey: bulk properties of previously optically identified sources.

J/A+A/281/375/  table1 table2 table3 table4  (17Kb)

FERETTI L, GIOVANNINI G. Astron. Astrophys. 281, 375 (1994) Structures of small size radio galaxies in clusters; Morphology

J/A+A/281/465/  annex.dat table1.dat table2.dat table3.dat table4.dat  (33Kb)

Andrievsky S.M., Kovtyukh V.V., Usenko I.A. Astron. Astrophys. 281, 465 (1994) The chemical composition of the s-Cepheids. I. Alpha Ursae Minoris (Polaris) and HR 7308 (V473 Lyr): unique Cepheids of the Galaxy

J/A+A/281/811/  table1 table2 table4  (18Kb)

IBANOGLU C., KESKIN V., AKAN M.C., EVREN S., TUNCA Z. Astron. Astrophys. 281, 811 (1994) Long-term luminosity variations and period changes in V471 Tau

J/A+A/281/817/ table1a table1b table2a table2b table3a table3b table4a table4b table5a table5b (45Kb) CASTELLI F., KURUCZ R.L. Astron. Astrophys. 281, 817 (1994) Model atmospheres for Vega

J/A+A/282/L9/  table1 table1.tex  (15Kb)

JETSU L., TUOMINEN I., GRANKIN K.N., MEL’NIKOV S.YU., SHEVCHENKO V.S. Astron. Astrophys. 282, L9 (1994) New ”flip–flop” of FK Comae Berenices

J/A+A/282/137/  table2.dat  (8Kb)

Annuk K. Astron. Astrophys. 282, 137 (1994) Photometry of the Wolf-Rayet binary HD211853 in the Vilnius photometric system

J/A+A/282/436/  refs.dat table1.dat table2.dat  (38Kb)

Maciel W.J., Koeppen J. Astron. Astrophys. 282, 436 (1994) Abundance gradients from disk planetary nebulae: O, Ne, S, and Ar

J/A+A/282/663 See VI/66/

J/A+A/282/831/  table4.dat table5.dat table6.dat  (10Kb)

Tokovinin A.A., Duquennoy A., Halbwachs J.-L., Mayor M. Astron. Astrophys. 282, 831 (1994) Duplicity in the solar neighbourhood. VII. Spectroscopic orbits of three K-dwarfs stars

J/A+A/282/899/ table1 table2 table3 table4 table5 table6 table7 table8 table9 table10 table11 table12 table13 table14 (36Kb) Blackwell D.E., Lynas-GRay A.E. Astron. Astrophys. 282, 899 (1994) Stellar effective temperatures and angular diameters determined by the infrared flux method (IRFM): revisions using improved Kurucz LTE stellar atmospheres

J/A+A/283/12/  table2 table3 table4  (33Kb)

BOSELLI A., GAVAZZI G. Astron. Astrophys. 283, 12 (1994) Multifrequency windows on spiral galaxies. III. Internal extinction at optical and near infrared wavelengths

J/A+A/283/319/  table table6  (293Kb)

Zhang H.L., Graziani M., Pradhan A.K. Astron. Astrophys. 283, 319 (1994) Atomic Data from the Iron Project III. Rate coefficients for electron impact excitation of Boron-like ions Ne VI, Mg VIII, Al IX, Si X, S XII, Ar XIV, Ca XVI and Fe XXII.

J/A+A/283/759/  qso.dat table2.dat table3.dat  (27Kb)

Petitjean P., Bergeron J. Astron. Astrophys. 283, 759 (1994) CIV QSO absorption systems and properties of galactic haloes at high redshift

J/A+A/283/779 See VII/184/

J/A+A/283/805/  table1 table3  (15Kb)

BICA E., ALLOIN D., SCHMITT H.R. Astron. Astrophys. 283, 805 (1994) Integrated spectral properties of star clusters in the the near-ultraviolet

J/A+A/283/911/  table1  (24Kb)

GRATTON R.G., CONTARINI G. Astron. Astrophys. 283, 911 (1994) Elemental abundances in the old open clusters NGC 2243 and Melotte 66

J/A+A/284/233/  table1  (43Kb)

CODELLA C., FELLI M., NATALE V. Astron. Astrophys. 284, 233 (1994) HII Regions and IRAS PSC sources: the reliability of the association

J/A+A/284/515/  table4 table4.tex  (37Kb)

TELTING J.H., KAPER L. Astron. Astrophys. 284, 515 (1994) Long-term periodic variability in UV absorption lines of the Be star gamma Cas: on the relation with V/R variations in the Hbeta line.

J/A+A/285/1/  table6  (10Kb)

COMTE G., AUGARDE R., CHALABAEV A., KUNTH D., MAEHARA H. Astron. Astrophys. 285, 1 (1994) Spectrographic study of a large sample of Kiso ultraviolet-excess galaxies. II - Discussion

J/A+A/285/247/ objects.dat table3.dat table4.dat table5.dat table6.dat table7.dat table8.dat (33Kb) Bujarrabal V., Fuente A., Omont A. Astron. Astrophys. 285, 247 (1994) Molecular observations of O- and C-rich circumstellar envelopes

J/A+A/285/404/  table1.dat  (15Kb)

Moles M., del Olmo A., Perea J., Masegosa J., Marquez I., Costa V. Astron. Astrophys. 285, 404 (1994) Star formation and merging in compact groups of galaxies

J/A+A/285/459/  table1 table2  (37Kb)

KREINER J.M., PAJDOSZ G., TREMKO J., ZOLA S. Astron. Astrophys. 285, 459 (1994) Investigation of the semidetached eclipsing binary RZ Draconis

J/A+A/285/473/  table1 table1.tex  (108Kb)

FORESTINI M. Astron. Astrophys. 285, 473 (1994) Low-mass stars: pre-main sequence evolution and nucleosynthesis; Abundances

J/A+A/285/785/  appena.dat  (15Kb)

Thimm G.J., Roeser H.-J., Hippelein H., Meisenheimer K. Astron. Astrophys. 285, 785 (1994) The galaxy population in the cluster Cl 1409+524 (3C 295)

J/A+A/285/812/  table1  (12Kb)

BRINKMANN W., SIEBERT J. Astron. Astrophys. 285, 812 (1994) Spectral properties of X-ray loud extragalactic radio sources

J/A+A/286/17/  table1 table2  (19Kb)

Seeberger R., Huchtmeier W.K., Weinberger R. Astron. Astrophys. 286, 17 (1994) 21cm observations of galaxies in the region of the Pisces-Perseus-Supercluster near the Galactic Plane

J/A+A/286/136/  table12  (9Kb)

AERTS C., WAELKENS C., DE PAUW M. Astron. Astrophys. 286, 136 (1994) Mode identification with the moment method in four multiperiodic beta Cephei stars: KK Velorum, nu Eri, Beta CMa, and V348 Normae

J/A+A/286/444/  table2 table3  (217Kb)

ORTOLANI S., BICA E., BARBUY B. Astron. Astrophys. 286, 444 (1994) The globular cluster NGC 6652

J/A+A/287/591/  table2  (8Kb)

JASNIEWICZ G., LAPIERRE G., MONIER R. Astron. Astrophys. 287, 591 (1994) Ultraviolet variations in the spectrum of the visible central star of Abell 35

J/A+A/287/769/  table2  (16Kb)

Burgarella D., Paresce F., Meylan G., King I.R., Greenfield P., Baxter D., Jedrzejewski R., Nota A., Albrecht R., Barbieri C., Blades J.C., Boksenberg A., Crane P., Deharveng J.M., Disney M.J., Jakobsen P., Kamperman T.M., Macchetto F., Mackay C.D., Weigelt G. Astron. Astrophys. 287, 769 (1994) HST observations of the core of globular cluster NGC 6397

J/A+A/287/817/  table2  (13Kb)

KESKIN V., IBANOGLU M.C., AKAN C., EVREN S., TUNCA Z. Astron. Astrophys. 287, 817 (1994) Luminosity variations and orbital period changes in RT Lacertae

J/A+A/287/927/  table3  (77Kb)

GRATTON R.G., SNEDEN C. Astron. Astrophys. 287, 927 (1994) Abundances of neutron-capture elements in metal-poor stars

J/A+A/287/990/  table3a table3b table3c  (23Kb)

Jenniskens P. Astron. Astrophys. 287, 990 (1994) Meteor stream activity. 1. The annual streams.

J/A+A/288/413/  table1.dat  (10Kb)

Deharveng J.-M., Albrecht R., Barbieri C., Blades J.C., Boksenberg A., Crane P., Disney M.J., Jakobsen P., Kamperman T.M., King I.R., Macchetto F., Mackay C.D., Paresce F., Weigelt G., Baxter D., Greenfield P., Jedrzejewski R., Nota A., Sparks W.B. Astron. Astrophys. 288, 413 (1994) The massive star content of the blue dwarf galaxy I Zw 36 from Faint Object Camera observations.

J/A+A/288/601/  table3 table4 table5 table6 table7 table8  (13Kb)

ZINCHENKO I., FORSSTROEM V., LAPINOV A., MATTILA K. Astron. Astrophys. 288, 601 (1994) Studies of dense molecular cores in regions of massive star formation. CS J=2-1 and HCN J=1-0 observations of 11 northern cores

J/A+A/288/713/  table1.dat table2.dat table3.dat  (56Kb)

Firmani C., Tutukov A.V. Astron. Astrophys. 288, 713 (1994) Bursting and stationary star formation in disks and nuclei of galaxies.

J/A+A/289/101/  table1a table1b table1c table1d  (90Kb)

BROWN A.G.A., DE GEUS E.J., DE ZEEUW P.T. Astron. Astrophys. 289, 101 (1994) The Orion OB1 association. I. Stellar content

J/A+A/289/137/  table  (45Kb)

AWADALLA N.S. Astron. Astrophys. 289, 137 (1994) Photoelectric light curve study of the DK Cygni system

J/A+A/289/740/  table2 table6 table7 table8 tables.tex  (80Kb)

KAUFER A., SZEIFERT T., KRENZIN R., BASCHEK B., WOLF B. Astron. Astrophys. 289, 740 (1994) The galactic abundance gradients traced by B-type stars

J/A+A/289/871/  aephe yyeri  (56Kb)

MACERONI C., VAN’T VEER F. Astron. Astrophys. 289, 871 (1994) Period variations of the late type contact binaries YY Eri and AE Phe: how to use light curves outside minimum

J/A+A/290/69/  table4 table6  (591Kb)

BUONANNO R., CORSI C.E., BUZZONI A., CACCIARI C., FERRARO F.R., FUSI PECCI F. Astron. Astrophys. 290, 69 (1994) The stellar population of the globular cluster M 3. I. Photographic photometry of 10 000 stars.

J/A+A/290/609/  table1  (16Kb)

DOUGHERTY S.M., WATERS L.B.F.M., BURKI G., COTE J., CRAMER N., VAN KERKWIJK M.H., TAYLOR A.R. Astron. Astrophys. 290, 609 (1994) Near-IR excess of Be stars

J/A+A/290/885/  table3 table3.tex  (60Kb)

PRIMAS F., MOLARO P., CASTELLI F. Astron. Astrophys. 290, 885 (1994) Abundances of four very metal-poor stars of the BPS survey

J/A+A/291/110/  table  (50Kb)

OLAH K., BUDDING E., KIM H.-I., ETZEL P.B. Astron. Astrophys. 291, 110 (1994) The active close binary system ER Vulpeculae

J/A+A/291/121/  table1 table2 table3  (20Kb)

STORM J., NORDSTROEM B., CARNEY B.W., ANDERSEN J. Astron. Astrophys. 291, 121 (1994) A Baade-Wesselink analysis of the RR Lyrae star V9 in 47 Tucanae.

J/A+A/291/155/  table1 table1.tex  (118Kb)

SALGADO M., BONAZZOLA S., GOURGOULHON E., HAENSEL P. Astron. Astrophys. 291, 155 (1994) High precision rotating neutron star models. I. Analysis of neutron star properties

J/A+A/291/261/  table1 table1.tex table2 table2.tex  (113Kb)

CODELLA C., FELLI M., NATALE V., PALAGI F., PALLA F. Astron. Astrophys. 291, 261 (1994) The occurrence of H2O masers in HII regions

J/A+A/291/943/  table1.dat  (28Kb)

Ossenkopf V., Henning T. Astron. Astrophys. 291, 943 (1994) Dust opacities for protostellar cores

J/A+A/293/75/  table3  (10Kb)

EDVARDSSON B., PETTERSSON B., KHARRAZI M., WESTERLUND B. Astron. Astrophys. 293, 75 (1995) Abundance analysis and origin of the zeta Sculptoris open cluster

J/A+A/293/347/  appendix  (55Kb)

HILL V., ANDRIEVSKY S., SPITE M. Astron. Astrophys. 293, 347 (1995) Chemical evolution of the Magellanic Clouds. VI. Chemical composition of nine F supergiants from different regions of the Large Magellanic Cloud

J/A+A/293/703/  table2 table3  (45Kb)

McKEITH C.D., GREVE A., DOWNES D., PRADA F. Astron. Astrophys. 293, 703 (1995) The outflow in the halo of M 82

J/A+A/293/833/  table1.dat  (9Kb)

David P., Pegourie B. Astron. Astrophys. 293, 833 (1995) The complex dielectric function for circumstellar silicate grains based on the IRAS data

J/A+A/293/953/  config.dat gamma.dat  (2.3 Mb)

Zhang H.L., Pradhan A.K. Astron. Astrophys. 293, 953 (1995) Atomic data from the IRON project: VI. Collision strengths and rate coefficients for Fe II,

J/A+A/293/967/  aij.dat aij2.dat ls.dat ls2.dat prog.f  (14.4 Mb)

Nahar S.N. Astron. Astrophys. 293, 967 (1995) Atomic data from the IRON Project. VII. Radiative transition probabilities for Fe II. (Rev. November 1995)

J/A+A/294/135/  table1 table3 table4 tables.tex  (40Kb)

STEFL S., BAADE D., HARMANEC P., BALONA L.A. Astron. Astrophys. 294, 135 (1995) Simultaneous photometric and spectroscopic monitoring of rapid variations of the Be star Eta Centauri

J/A+A/294/377/  table2 table3 table4 table5 table6 table7 table8 tables.tex  (107Kb)

VOGEL S., REIMERS D. Astron. Astrophys. 294, 377 (1995) The ultraviolet absorption spectrum of the z=2.72 QSO HS 1700+6416. I. Results on heavy-element absorption systems.

J/A+A/295/101/  table4  (66Kb)

Jahn K., Kaluzny J., Rucinski S.M. Astron. Astrophys. 295, 101 (1995) Variable stars in the field of the open cluster NGC 7789

J/A+A/295/206/  table1a.dat table1b.dat table1c.dat  (23Kb)

Jenniskens P. Astron. Astrophys. 295, 206 (1995) Meteor stream activity. II. Meteor outbursts

J/A+A/295/585/  table1  (14Kb)

Sandqvist A., Jorsater S., Lindblad P.O Astron. Astrophys. 295, 585 (1995) The central region of NGC 1365. VLA and SEST observations of the radio continuum and CO.

J/A+A/296/319/  tables  (18Kb)

D’ONOFRIO M., ZAGGIA S., LONGO G., CAON N., CAPACCIOLI M. Astron. Astrophys. 296, 319 (1995) Major axis kinematics of 15 early-type galaxies in the Fornax cluster.

J/A+A/296/643/  table1 table2  (16Kb)

Karachentsev I.D., Karachentseva V.E., Richter G.M., Vennik J.A. Astron. Astrophys. 296, 643 (1995) A search for low surface brightness dwarf galaxies in the Coma cluster core

J/A+A/296/665/  table3  (16Kb)

VERON P., HAWKINS M.R.S. Astron. Astrophys. 296, 665 (1995) A survey for medium redshift (2.2 z 3.0) optically variable QSOs.

J/A+A/296/680/  table2 table3  (1.2 Mb)

ORTOLANI S., BICA E., BARBUY B. Astron. Astrophys. 296, 680 (1995) Two reddened globular clusters projected close to the Galactic center: Palomar 6 and Djorgovski 1

J/A+A/296/727/  table1 table2 table3 table4 tables.tex  (56Kb)

SZYMCZAK M., ENGELS D. Astron. Astrophys. 296, 727 (1995) H2O maser emission from semiregular variables

J/A+A/297/28/  table3  (14Kb)

BONFANTI P., RAMPAZZO R., COMBES F., PRUGNIEL P., SULENTIC J.W. Astron. Astrophys. 297, 28 (1995) Stellar dynamics in E+E pairs of galaxies. I. NGC 741/742, 1587/88 and 2672/73. The data.

J/A+A/297/49/  table1 table2  (13Kb)

PETROSIAN A.R., TURATTO M. Astron. Astrophys. 297, 49 (1995) The spatial distribution of supernovae in paired and interacting galaxies

J/A+A/297/61/  table8 table9  (12Kb)

BELLONI P., BRUZUAL A.G., THIMM G.J.,ROESER H.-J. Astron. Astrophys. 297, 61 (1995) Detectability and incidence of E+A galaxies in the distant cluster Cl0939+472 (z=0.41)

J/A+A/297/359/  table1  (8Kb)

Parthasarathy M., Sarma M.B.K., Vivekananda Rao P. Astron. Astrophys. 297, 359 (1995) Photometric elements, absolute dimensions and evolutionary status of the eclipsing binary HU Tauri (HR 1471)

J/A+A/297/391/  refs.dat table1 table2 table3 table4 table5 table6 tablea1  (81Kb)

Neuhaeuser R., Sterzik M.F., Schmitt J.H.M.M., Wichmann R., Krautter J. Astron. Astrophys. 297, 391 (1995) ROSAT Survey observation of T Tauri stars in Taurus

J/A+A/297/494/  table1 table2  (17Kb)

SZYMCZAK M., LE SQUEREN A.M., P.SIVAGNANAM, TRAN MINH F., FOURNIER A. Astron. Astrophys. 297, 494 (1995) A survey of main-line OH maser emission from semiregular variables

J/A+A/297/503/  table2  (15Kb)

CASTILHO B.V., BARBUY B., GREGORIO-HETEM J. Astron. Astrophys. 297, 503 (1995) Analysis of the moderately Li-rich giant HD 146850

J/A+A/297/617/  table1.dat table2.dat table3.dat  (30Kb)

Kraan-Korteweg R.C., Fairall A.P., Balkowski C. Astron. Astrophys. 297, 617 (1995) Extragalactic large-scale structures behind the southern Milky Way. I. Redshifts obtained at the SAAO in the Hydra/Antlia extension.

J/A+A/298/115/  table2 table2.tex table4 table4.tex table5 table5.tex  (35Kb)

Randich S., Schmitt J.H.M.M. Astron. Astrophys. 298, 115 (1995) A ROSAT X-ray study of the Praesepe cluster

J/A+A/298/375/  table1 table2  (21Kb)

FALCKE H., MALKAN M.A., BIERMANN P.L. Astron. Astrophys. 298, 375 (1995) The jet-disk symbiosis II. Interpreting the Radio/UV correlations in quasars

J/A+A/298/461/  table.dat  (34Kb)

Ferraro F.R., Fusi Pecci F., Montegriffo P., Origlia L., Testa V. Astron. Astrophys. 298, 461 (1995) Deep infrared array photometry of galactic globular clusters: the main sequence of NGC 6171

J/A+A/299/39/  table1.dat table2.dat  (21Kb)

Geier S., Wendker H.J., Wisotzki L. Astron. Astrophys. 299, 39 (1995) ROSAT PSPC observations of the Orion Trapezium area. I. Pre-main sequence and O stars.

J/A+A/299/53/  i4651 n3680  (18Kb)

MERMILLIOD J.-C., ANDERSEN J., NORDSTROM B., MAYOR M. Astron. Astrophys. 299, 53 (1995) Red giants in open clusters: IV. NGC 3680 and IC 4651

J/A+A/299/89/  table1  (25Kb)

BOUVIER J., COVINO E., KOVO O., MARTIN E.L., MATTHEWS J.M., TERRANEGRA L., BECK S.C. Astron. Astrophys. 299, 89 (1995) COYOTES II: Spot properties and the origin of photometric period variations in T Tauri stars

J/A+A/299/238/  table1 table3  (14Kb)

VAN DE STEENE G.C., POTTASCH S.R. Astron. Astrophys. 299, 238 (1995) Radio continuum observations of planetary nebula candidates from the northern hemisphere

J/A+A/299/347/  table2 table2.tex  (13Kb)

CHAMARAUX P., KAZES I., SAITO M., YAMADA T., TAKATA T. Astron. Astrophys. 299, 347 (1995) A search for IRAS galaxies behind the Taurus molecular clouds.

J/A+A/299/453/  table1 table2  (46Kb)

LEPINE J.R.D., ORTIZ R., EPCHTEIN N. Astron. Astrophys. 299, 453 (1995) OH/IR stars: near-infrared photometry, and discussion of the Mira-OH/IR sequence

J/A+A/299/557/  table  (21Kb)

LECAVELIER DES ETANGS A., DELEUILE M., VIDAL-MADJAR A., FERLET R. NITSCHLEM C., NICOLET B., LAGRANGE-HENRI A.M. Astron. Astrophys. 299, 557 (1995) Beta Pictoris: Evidence of light variations

J/A+A/299/621/  appen table2 table3 tables.tex  (69Kb)

BENEST D., DUVENT J.L. Astron. Astrophys. 299, 621 (1995) Is Sirius a triple star?

J/A+A/299/696/  table1 table1.tex  (86Kb)

SCHILBACH E., ROBICHON N., SOUCHAY J., GUIBERT J. Astron. Astrophys. 299, 696 (1995) Membership probabilities in the Pleiades field

J/A+A/299/755/  table3 table4 table5  (37Kb)

BLOECKER T. Astron. Astrophys. 299, 755 (1995) Stellar evolution of low and intermediate-mass-stars. II. Post-AGB evolution.

J/A+A/300/128/  eso92 eso96  (66Kb)

CARRARO G., VALLENARI A., ORTOLANI S. Astron. Astrophys. 300, 128 (1995) CCD photometry of the faint old open clusters ESO 96-SC04 and ESO 92-SC18

J/A+A/300/323/  table1  (19Kb)

JAUNSEN A.O., JABLONSKI M., PETTERSEN B.R., STABELL R. Astron. Astrophys. 300, 323 (1995) The NOT gravitational lens survey for multiply imaged quasars

J/A+A/300/726/  table2 table3  (746Kb)

ORTOLANI S., BICA E., BARBUY B. Astron. Astrophys. 300, 726 (1995) BH 176 and AM-2: globular or open clusters ?

J/A+A/300/751/  table1 table3 table4 table5  (58Kb)

AXER M., FUHRMANN K., GEHREN T. Astron. Astrophys. 300, 751 (1995) Spectroscopic analyses of metal-poor stars. II. The evolutionary stage of subdwarfs.

J/A+A/301/396/  table3  (138Kb)

WILL J.M., VAZQUEZ A., FEINSTEIN A., SEGEEWISS W. Astron. Astrophys. 301, 396 (1995) CCD Photometry of the Young Association NGC 1962-65-66-70 in the Large Magellanic Cloud

J/A+A/303/137/  table1.dat  (17Kb)

Beaulieu J.P., Grison P., Tobin W., Pritchard J.D., Ferlet R., Lepeintre F., Vidal-Madjar A., Maurice E., Prevot L., Gry C., Guibert J., Moreau O., Tajhamady F., Aubourg E., Bareyre P., Coutures C., Gros M., Laurent B., Lachieze-Rey M., Lesquoy E., Magneville C., Milsztajn A., Moscoso L., Queinnec F., Renault C., Rich J., Spiro M., Vigroux L., Zylberajch S., Ansari R., Cavalier F., Moniez M. Astron. Astrophys. 303, 137 (1995) EROS variable stars: fundamental-mode and first-overtone Cepheids in the bar of the Large Magellanic Cloud.

J/A+A/303/175/  table2 table2.tex  (123Kb)

BALUTEAU J.-P., ZAVAGNO A., MORISSET C., PEQUIGNOT D. Astron. Astrophys. 303, 175 (1995) The 654-1046nm line spectrum of the planetary nebula NGC 7027

J/A+A/303/420/  table1 table2  (27Kb)

Chyzy K.T., Zieba S. Astron. Astrophys. 303, 420 (1995) Cosmological evolution and large scale structures of radio galaxies and quasars

J/A+A/303/468/  table3 table3.tex  (27Kb)

Minniti D. Astron. Astrophys. 303, 468 (1995) Spectroscopy and IR photometry for giant stars in obscured globular clusters: NGC 6325, NGC 6401, NGC 6517, NGC 6642, HP 1 and Pal 6.

J/A+A/303/661/  table1 table2  (25Kb)

Donas J., Milliard B., Laget M. Astron. Astrophys. 303, 661 (1995) Ultraviolet observations of galaxies in nearby clusters. III. Star-forming galaxies in the Coma Cluster

J/A+A/303/893/  table1 table2  (22Kb)

GORNY S.K., STASINSKA G. Astron. Astrophys. 303, 893 (1995) On the status of planetary nebulae with WR-type nuclei

J/A+A/304/235/  table6a table6b  (53Kb)

Bozic H., Harmanec P., Horn J., Koubsky P., Scholz G., McDavid D., Hubert A.-M., Hubert H. Astron. Astrophys. 304, 235 (1995)

Toward a consistent model of the B0.5IVe + sdO binary phi Per

J/A+A/304/347/ fig1a.ps fig1b.ps fig1c.ps fig1d.ps h13.dat h15.dat h21.dat h30.dat h36.dat h37.dat h39.dat h44.dat h47.dat h51.dat h52.dat h58.dat h67.dat h71.dat h78.dat h90.dat h91.dat h92.dat h96.dat h100.dat p5.dat p10.dat p12.dat p27.dat p28.dat p43.dat p47.dat p48.dat p56.dat p62.dat p89.dat p102.dat p106.dat p118.dat p119.dat p122.dat p123.dat p126.dat p131.dat p134.dat stars (1.4 Mb) Meliani M.T., Barbuy B., Richtler T. Astron. Astrophys. 304, 347 (1995) Metallicities and carbon abundances of 40 field red supergiants in the Small Magellanic Cloud

J/A+A/304/406/  photo velo  (67Kb)

Chapellier E., Le Contel J.M., Le Contel D., Sareyan J.P., Valtier J.C. Astron. Astrophys. 304, 406 (1995) Pulsation and binarity in Beta Cephei Stars. II. 16 Lac

J/A+A/304/415/  table1  (19Kb)

van Genderen A.M., Sterken C., de Groot M., Stahl O., Andersen J., Andersen M.I., Caldwell J.A.R., Casey B., Clement R., Corradi W.J.B., Cuypers J., Debehogne H., Garcia de Maria J.M., Joench-Soerensen H., Vaz L.P.R., Stefl S., Suso Lopez J., Beele D., Eggenkamp I.M.M.G., Goecking K.-D., Jorissen A., de Koff S., Kuss C., Schoenmakers A.P., Vink J., Waelde E. Astron. Astrophys. 304, 415 (1995) A pulsating star inside Eta Carinae I. Light variations 1992-1994

J/A+A/304/563/  table1  (25Kb)

Phillips K.J.H., Feldman U. Astron. Astrophys. 304, 563 (1995) Properties of Cool Flares with GOES class B5 to C2

J/A+A/305/74/  table3 table6 table7 table17 table18 table19 table20  (34Kb)

Vogler A., Pietsch W., Kahabka P. Astron. Astrophys. 305, 74 (1996) ROSAT observations of three edge-on spiral galaxies

J/A+A/305/245/  table4 table4.tex  (21Kb)

Magain P., Zhao G. Astron. Astrophys. 305, 245 (1996) Empirical study of departures from the excitation equilibrium of Fe I in metal-poor stars

J/A+A/305/551/ table10 table11 table12 table13 table14 table15 table16 table17 table18 tables.tex (38Kb) Andrievsky S.M., Kovtyukh V.V, Usenko I.A. Astron. Astrophys. 305, 551 (1996) The chemical composition of the s-Cepheids. II

J/A+A/305/960/  table1  (13Kb)

Chin Y.-N., Henkel C., Whiteoak J.B., Langer N., Churchwell E.B. Astron. Astrophys. 305, 960 (1996) Interstellar sulfur isotopes and stellar oxygen burning

J/A+A/305/989 See I/219/

J/A+A/306/9/  tab3-5.dat table2.dat  (19Kb)

Soffner T., Mendez R.H., Jacoby G.H., Ciardullo R., Roth M.M., Kudritzki R.P. Astron. Astrophys. 306, 9 (1996) Planetary nebulae and HII regions in NGC 300.

J/A+A/306/134/  table2  (205Kb)

Ortolani S., Bica E., Barbuy B. Astron. Astrophys. 306, 134 (1996) NTT V, I and Gunn z colour-magnitude diagrams of Liller 1: a globular cluster as metal-rich as the inner bulge stellar population ?

J/A+A/306/924/  fig9.ps table11  (526Kb)

Bakker E.J., Van Der Wolf F.L.A., Lamers H.J.G.L.M., Gulliver A.F., Ferlet R., Vidal-Madjar A. Astron. Astrophys. 306, 924 (1996) The optical spectrum of HR 4049 (includes a line identification from 3650 to 10850 Angstroems)

J/A+A/307/33/  table1 table2 table3  (33Kb)

Schartel N., Walter R., Fink H.H., Truemper J. Astron. Astrophys. 307, 33 (1996) Redshift dependence of soft X-ray quasar spectra

J/A+A/307/80/  n2099.dat  (23Kb)

Mermilliod J.-C., Huestamendia G., del Rio G., Mayor M. Astron. Astrophys. 307, 80 (1996) Red giants in open clusters: V. NGC 2099

J/A+A/307/215/  table7  (22Kb)

Cuisinier F., Acker A., Koeppen J. Astron. Astrophys. 307, 215 (1996) Spectrophotometric observations of planetary nebulae high above the Galactic plane

J/A+A/307/869/  table14 table15  (40Kb)

Bakker E.J., Lamers H.J.G.L.M., Waters L.B.F.M., Waelkens C., Trams N.R., van Winckel H. Astron. Astrophys. 307, 869 (1996) A study on the nature of the peculiar supergiant HD 101584

J/A+A/308/481/  table1  (9Kb)

Chadid M., Gillet D. Astron. Astrophys. 308, 481 (1996) Observation of the metallic line doubling phenomenon in the variable star RR Lyrae.

J/A+A/308/573/  t.datable2 table1.dat  (32Kb)

Molinari S., Brand J., Cesaroni R., Palla F. Astron. Astrophys. 308, 573 (1996) A search for precursors of ultracompact HII regions in a sample of luminous IRAS sources. I. Association with ammonia cores.

J/A+A/308/733/  table2  (185Kb)

Ortolani S., Barbuy B., Bica E. Astron. Astrophys. 308, 733 (1996) NTT VI photometry of the metal-rich and obscured bulge globular cluster Terzan 5.

J/A+A/309/81/  table1 table1.tex table2 table2.tex table3 table3.tex table4 table4.tex  (47Kb)

Wang T., Brinkmann W., Bergeron J. Astron. Astrophys. 309, 81 (1996) X-ray properties of active galactic nuclei with optical FeII emission

J/A+A/309/116/  table2.dat  (14Kb)

Johnston H.M., Verbunt F., Hasinger G. Astron. Astrophys. 309, 116 (1996) ROSAT observations of ten globular clusters with large core radii

J/A+A/309/446/  table2  (40Kb)

Petit H., Hua C.T., Bersier D., Courtes G. Astron. Astrophys. 309, 446 (1996) The spiral structure of M51 from Halpha and 2000A UV Images. A new tracer of density waves effects.

J/A+A/309/521/  table1  (8Kb)

Horn J., Kubat J., Harmanec P., Koubsky P., Hadrava P., Simon V., Stefl S., Skoda P. Astron. Astrophys. 309, 521 (1996) Spectroscopic orbit of the triple star 55 Ursae Majoris.

J/A+A/309/581/  table2.dat table3.dat table4.dat table5.dat  (34Kb)

Weikard H., Wouterloot J.G.A., Castets A., Winnewisser G., Sugitani K. Astron. Astrophys. 309, 581 (1996) The structure of the IC 1396 region.

J/A+A/309/655/  table1  (14Kb)

Shcherbakov A.G., Shcherbakova Z.A., Tuominen I., Jetsu L. Astron. Astrophys. 309, 655 (1996) He I lambda 10830A line as an indicator of solar chromospheric activity

J/A+A/309/749/  tablea1.dat tablea2.dat tablea3.dat tableb1.dat  (402Kb)

Prugniel P., Simien F. Astron. Astrophys. 309, 749 (1996) The fundamental plane of early-type galaxies: stellar populations and mass-to-light ratio

J/A+A/309/787/  table1 table2 table3 table4 table5 table8 tables.tex  (66Kb)

Stefl S., Balona L.A. Astron. Astrophys. 309, 787 (1996) Rapid photometric and spectroscopic variability of the Be star DX Eridani

J/A+A/310/8/  table6  (15Kb)

Katgert P., Mazure A., Perea J., den Hartog R., Moles M., Le Fevre O., Dubath P., Focardi P., Rhee G., Jones B., Escalera E., Biviano A., Gerbal D., Giuricin G. Astron. Astrophys. 310, 8 (1996) The ESO Nearby Abell Cluster Survey. I. Description of the dataset and definition of physical systems

J/A+A/310/31/  table1a table1b  (14Kb)

Mazure A., Katgert P., den Hartog R., Biviano A., Dubath P., Escalera E., Focardi P., Gerbal D., Giuricin G., Jones B., Le Fevre O., Moles M., Perea J., Rhee G. Astron. Astrophys. 310, 31 (1996) The ESO Nearby Abell Cluster Survey (ENACS). II. The distribution of velocity dispersions of rich galaxy clusters.

J/A+A/310/143/  tables  (7Kb)

Fernandez M., Eiroa C. Astron. Astrophys. 310, 143 (1996) Variability of classical T Tauri stars. Its relation to the accretion process

J/A+A/310/164/  table1  (11Kb)

De Mey K., Aerts C., Waelkens C., van Winckel H. Astron. Astrophys. 310, 164 (1996) The early-type multiple system Eta Orionis. II. Line profile variations in component Ab

J/A+A/310/173/  table1.dat table1.tex  (20Kb)

Neff J.E., Pagano I., Rodono M., Brown A., Dempsey R.C., Fox D.C., Linsky J.L. Astron. Astrophys. 310, 173 (1996) Rotational modulation and flares on RS Canum Venaticorum and BY Draconis stars. XIX. Simultaneous IUE, ROSAT, VLA, and visual observations of TY Pyxidis

J/A+A/310/353/  table2 table8  (14Kb)

Guthrie B.N.G., Napier W.M. Astron. Astrophys. 310, 353 (1996) Redshift periodicity in the Local Supercluster

J/A+A/310/384/  table  (28Kb)

Greiner J., Danner R., Bade N., Richter G.A., Kroll P., Komossa S. Astron. Astrophys. 310, 384 (1996) Four new active galaxies with steep soft X-ray spectra

J/A+A/310/715/  tucana.dat  (37Kb)

Castellani M., Marconi G., Buonanno R. Astron. Astrophys. 310, 715 (1996) The Tucana dwarf Galaxy.

J/A+A/310/879/  table1 table1.tex  (28Kb)

Barrado y Navascues D., Stauffer J.R. Astron. Astrophys. 310, 879 (1996) Lithium abundance in binaries of the Hyades open cluster

J/A+A/310/893/  table13 table14 table15 table16 table17 table18  (30Kb)

Bakker E.J., Waters L.B.F.M., Lamers H.J.G.L.M., Trams N.R., van der Wolf F.L.A. Astron. Astrophys. 310, 893 (1996) Detection of C_{2}, CN, and NaID absorption in the AGB remnant of HD 56126

J/A+A/310/933/  table1 table1.tex table2 table2.tex  (32Kb)

Ohnaka K., Tsuji T. Astron. Astrophys. 310, 933 (1996) Quantitative analysis of carbon isotopic ratios in carbon stars. I. 62 N-type and 15 SC-type carbon stars.

J/A+A/311/211/  table2  (13Kb)

Kuerster M., Schmitt J.H.M.M. Astron. Astrophys. 311, 211 (1996) Forty days in the life of CF Tucana (=HD 5303). The longest stellar X-ray flare observed with ROSAT

J/A+A/311/361/  table2  (39Kb)

Tantalo R., Chiosi C., Bressan A., Fagotto F. Astron. Astrophys. 311, 361 (1996) Spectro-photometric evolution of elliptical galaxies. II. Models with infall.

J/A+A/311/384/  table2 table3  (15Kb)

Lamer G., Brunner H., Staubert R. Astron. Astrophys. 311, 384 (1996) ROSAT observations of BL Lacertae objects.

J/A+A/311/484/  table1 table1.tex table2 table2.tex  (16Kb)

Munari U., Yudin B.F., Kolotilov E.A., Tomov T.V. Astron. Astrophys. 311, 484 (1996) UBV-JHKLM photometry of CH Cygni over 1978-1995: dust properties and doubts on the triple star model.

J/A+A/311/523/  table1  (9Kb)

Maceroni C., van’t Veer F. Astron. Astrophys. 311, 523 (1996) The properties of W Ursae Majoris contact binaries: new results and old problems

J/A+A/311/710/ barycent.xyz code emb.xyz faintsat.zip install.txt j6.xyz j7.xyz j8.xyz j9.xyz jupiter.xyz satxjup.f table3 table4 table5 table6 (3.2 Mb) Rocher P., Chapront J. Astron. Astrophys. 311, 710 (1996) Observations and ephemerides of the faint satellites of Jupiter

J/A+A/311/873/  table  (39Kb)

Waelkens C. Astron. Astrophys. 311, 873 (1996) The slowly pulsating B stars HD 34798 and HD 45284.

J/A+A/312/74/  table3  (54Kb)

Tucholke H.-J., Scholtz R.-D., Brosche P. Astron. Astrophys. 312, 74 (1996) Proper motion study of the globular cluster M92.

J/A+A/312/93/  refs table1 table1.tex  (38Kb)

Van Paradijs J., Augusteijn T., Stehle R. Astron. Astrophys. 312, 93 (1996) The systemic-velocity distribution of cataclysmic variables

J/A+A/312/111/  table1  (30Kb)

Jurcsik J., Kovacs G. Astron. Astrophys. 312, 111 (1996) Determination of [Fe/H] from the light curves of RR Lyrae stars.

J/A+A/312/221/  table1 table2 table3 table4  (33Kb)

Montes D., Fernandez-Figueroa M.J., Cornide M., De Castro E. Astron. Astrophys. 312, 221 (1996) The behaviour of the excess CaII H + K and Hepsilon emissions in chromospherically active binaries

J/A+A/312/439/ figure5a.ps figure5b.ps figure5c.ps figure5d.ps figure5e.ps figure5f.ps figure5g.ps figure5h.ps table4 table5 table6 table7 tables.tex (2.1 Mb) Wichmann R., Krautter J., Schmitt J.H.M.M., Neuhauser R., Alcala J.M., Zinnecker H., Wagner R.M., Mundt R., Sterzik M.F. Astron. Astrophys. 312, 439 (1996) New weak-line T Tauri stars in Taurus-Auriga

J/A+A/312/745/  table2 table2.tex table3 table3.tex  (33Kb)

Ramella M., Focardi P., Geller M.J. Astron. Astrophys. 312, 745 (1996) The redshift-space neighbourhoods of 13 SSRS groups of galaxies

J/A+A/312/818/  table2 table2.tex table3 table3.tex table6 table6.tex  (50Kb)

Dachs J., Hummel W. Astron. Astrophys. 312, 818 (1996) ROSAT survey of stellar X-ray sources in the young open cluster NGC 2516

J/A+A/312/879/  note4 table2a table2b table2c table2d table4  (346Kb)

Harmanec P., Morand F., Bonneau D., Jiang Y., Yang S., Guinan E.F., Hall D.S., Mourard D., Hadrava P., Bozic H., Sterken C., Tallon-Bosc I., Walker G.A.H., McCook G.P., Vakili F., Stee P., Le Contel J.M. Astron. Astrophys. 312, 879 (1996) Jet-like structures in beta Lyrae. Results of optical interferometry, spectroscopy and photometry

J/A+A/312/1017/ table41 table42 table51 table51.tex table52 table52.tex table61 table61.tex table62 table62.tex (37Kb) Souchay J., Kinoshita H. Astron. Astrophys. 312, 1017 (1996) Corrections and new developments in rigid Earth nutation theory. I. Lunisolar influence including indirect planetary effects.

J/A+A/312/1031/  table2 table3  (13Kb)

Argyle R.W., Einicke O.H., Pilkington J.D.H., Morrison L.V., Fabricius C., Helmer L. Astron. Astrophys. 312, 1031 (1996) Comparison of the Carlsberg optical references frame with the International celestial reference frame

J/A+A/313/129/  table1 table2  (11Kb)

Burgarella D., Buat V. Astron. Astrophys. 313, 129 (1996) Stellar populations and color gradients in the post core collapse globular cluster M30

J/A+A/313/377/  appen.tex bivk1.ps bivk2.ps bvrk1.ps bvrk2.ps coorfig2.ps  (294Kb)

De Jong R.S. Astron. Astrophys. 313, 377 (1996) Near-infrared and optical broadband surface photometry of 86 face-on disk dominated galaxies. IV. Using color profiles to study stellar and dust content of galaxies.

J/A+A/313/841/  appen.tex fig5a.ps fig5b.ps  (436Kb)

Dominik M., Hirshfeld A.C. Astron. Astrophys. 313, 841 (1996) Evidence for a binary lens in the MACHO LMC No.1 microlensing event

J/A+A/313/924/  table2 table2.tex  (20Kb)

Costa R.D.D., de Freitas Pacheco J.A., de Franca J.A.Jr Astron. Astrophys. 313, 924 (1996) Abundances in type I planetary nebulae: is the galactic disk presently oxygen deficient?

J/A+A/313/949/  table5 table5.tex table6 table6.tex  (33Kb)

Bocchialini K., Gouttebroze P. Astron. Astrophys. 313, 949 (1996) Solar chromospheric structures as observed simultaneously in strong UV lines. II. Network and cell modelling.

J/A+A/313/1008/  appen.tex  (29Kb)

Achterberg A. Astron. Astrophys. 313, 1008 (1996) Variational principle for slender flux tubes. I. General equations and added mass effects.

J/A+A/314/43/  table1 table1.tex  (30Kb)

Kollatschny W., Dietrich M. Astron. Astrophys. 314, 43 (1996) Broad emission-line profile variability in NGC 5548. The optical emission lines in 1989.

J/A+A/314/191/  tables  (24Kb)

Gratton R.G., Carretta E., Castelli F. Astron. Astrophys. 314, 191 (1996) Abundances of light elements in metal-poor stars. I. Atmospheric parameters and a new Teff scale.

J/A+A/314/312/  table1-9 tables.tex  (57Kb)

Arlot J.-E. Astron. Astrophys. 314, 312 (1996) Predictions of the mutual events of the Galilean satellites of Jupiter occurring in 1997

J/A+A/314/448/  table2 table3  (24Kb)

Lapasset E., Ahumada J. Astron. Astrophys. 314, 448 (1996) Discovery of an eclipsing blue straggler in NGC 2354

J/A+A/314/491/  table1 table2 table6 tables.tex  (23Kb)

Wade G., North P., Mathys G., Hubrig S. Astron. Astrophys. 314, 491 (1996) An analysis of the Ap spectroscopic binary HD 59435

J/A+A/314/541/  table1-2 table3-5 tables.tex  (47Kb)

Antonello E., Morelli P.L. Astron. Astrophys. 314, 541 (1996) Search for resonance effects in long period Cepheids

J/A+A/314/585/  appen.tex  (26Kb)

Guilain C., Mauron N. Astron. Astrophys. 314, 585 (1996) NaI/KI scattering observations in circumstellar envelopes: constraints on ionization and mass-loss rates.

J/A+A/314/615/ figure1.ps figure3.ps figure5.ps figure7.ps figure8.ps figure9.ps figure10.ps figure11.ps table1 table1.tex table2 table3 table4 (1.1 Mb) Liechti S., Wilson T.L. Astron. Astrophys. 314, 615 (1996) Maps of the 36 GHz methanol emission

J/A+A/314/846/  table2a table2b  (9Kb)

Duquennoy A., Tokovinin A.A., Leinert C., Glindemann A., Halbwachs J.-L. Mayor M. Astron. Astrophys. 314, 846 (1996) Spectroscopic and speckle interferometric orbit of Gliese 692.1

J/A+A/314/871/  vhyafig1.ps vhyafig2.ps vhyafig3.ps  (942Kb)

Kahane C., Audinos P., Barnbaum C., Morris M. Astron. Astrophys. 314, 871 (1996) V Hydrae : the missing link between spherical red giants and bipolar planetary nebulae? Radio observations of the molecular envelope.

J/A+A/314/896/ table1 table1.tex table2a table2a.tex table2b table2b.tex table3 table3.tex table6 table6.tex table7 table7.tex table8 table8.tex (62Kb) Le Sidaner P., Le Bertre T. Astron. Astrophys. 314, 896 (1996) Optical and infrared observations of 27 oxygen-rich stars. Modelling of the circumstellar dust shells.

J/A+A/315/40/  tuc  (21Kb)

Saviane I., Held E.V., Piotto G Astron. Astrophys. 315, 40 (1996) CCD photometry of the Tucana dwarf galaxy

J/A+A/315/95/  table2  (11Kb)

Rolleston W.R.J., Brown P.J.F., Dufton P.L., Howarth I.D Astron. Astrophys. 315, 95 (1996) The chemical compositions of three main-sequence B-type stars in the Large Magellanic Cloud

J/A+A/315/348/  tablea1.dat tablea2.dat tableph.tex  (24Kb)

Karachentsev I., Musella I. Astron. Astrophys. 315, 348 (1996) Two dwarf galaxies in Orion with low radial velocities.

J/A+A/315/384/  table3  (20Kb)

Raboud D. Astron. Astrophys. 315, 384 (1996) Binarity among B-stars in NGC 6231

J/A+A/315/578/  appen.tex  (21Kb)

Oliver R.J., Masheder M.R.W., Thaddeus P. Astron. Astrophys. 315, 578 (1996) A new CO Survey of the Monoceros OB1 region

J/A+A/316/102/  table1  (16Kb)

Tapia M., Persi P., Roth M. Astron. Astrophys. 316, 102 (1996) The embedded stellar population in northern NGC 6334

J/A+A/316/123/  appen.tex fig1.ps fig2.ps fig3.ps fig4.ps  (231Kb)

Lindner U., Fritze-v. Alvensleben U., Fricke K.J. Astron. Astrophys. 316, 123 (1996) Optically identified QSO absorption systems and galaxy evolution

J/A+A/316/147/  table1 table1.tex table2 table2.tex  (57Kb)

Fleming T.A., Snowden S.L., Pfeffermann E., Briel U., Greiner J. Astron. Astrophys. 316, 147 (1996) Catalogue and luminosity function of white dwarfs detected in the ROSAT All-Sky Survey

J/A+A/317/54/  table1 table2 tables.tex  (17Kb)

Jorgensen U.G., Jimemez R. Astron. Astrophys. 317, 54 (1997) Globular clusters 1 and 3 in the Fornax dwarf galaxy

J/A+A/317/328/  table5 table6 table7 table8 table9 tables.tex  (203Kb)

Supper R., Hasinger G., Pietsch W., Truemper J., Jain A., Magnier E.A. Lewin W.H.G., van Paradijs J. Astron. Astrophys. 317, 328 (1997) ROSAT PSPC survey of M 31

J/A+A/317/416/  table1 table1.tex  (36Kb)

Lespine Y., Petitjean P. Astron. Astrophys. 317, 416 (1997) Physical conditions in metal line systems toward Q1037-2704: Evidence for superclustering at z 2.

J/A+A/317/423/  table3.dat  (8Kb)

Patat F., Barbon R., Cappellaro E., Turatto M. Astron. Astrophys. 317, 423 (1997) Revised photometry and color distribution of Type Ia supernovae observed at Asiago in the seventies.

J/A+A/317/448/  table2  (11Kb)

Grebel E.K. Astron. Astrophys. 317, 448 (1997) Be star surveys with CCD photometry. II. NGC 1818 and its neighbouring cluster in the LMC.

J/A+A/317/503/  appen.tex fig11.ps table1 table2 table3 tables.tex  (40Kb)

Groenewegen M.A.T Astron. Astrophys. 317, 503 (1997) IRC+10 216 revisited I: the circumstellar dust shell

J/A+A/317/689/  sd.dat tables.tex wd.dat  (39Kb)

Thejll P., Flynn C., Williamson R., Saffer R. Astron. Astrophys. 317, 689 (1997) Proper motion of the hot subdwarfs. The kinematic population membership of the SdB

J/A+A/317/712/ table2 table2.tex table3a table3a.tex table3b table3b.tex table4 table4.tex table5 table5.tex (41Kb) Altamore A., Cassatella A. Astron. Astrophys. 317, 712 (1997) The 1978-1995 variability of the symbiotic star AG Pegasi in the ultraviolet

J/A+A/317/761/  table2  (13Kb)

Alvarez R., Mennessier M.-O. Astron. Astrophys. 317, 761 (1997) Determination of Miras temperatures from TiO and VO bands. Estimates of distances.

J/A+A/318/53/  table1 table1.tex  (23Kb)

Skopal A. Astron. Astrophys. 318, 53 (1997) Is the symbiotic binary EG And an eclipsing system?

J/A+A/318/111/  supertab.sty table11 table12 tables.tex  (49Kb)

Motch C., Guillout P., Haberl F., Pakull M.W., Pietsch W., Reinsch K. Astron. Astrophys. 318, 111 (1997) The ROSAT Galactic Plane Survey: analysis of a low latitude sample area in Cygnus.

J/A+A/318/256/  table1 table3  (23Kb)

Gorny S.K., Stasinska G., Tylenda R. Astron. Astrophys. 318, 256 (1997) Planetary nebulae morphologies, central star masses and nebular properties

J/A+A/318/347/  table2 table2.tex table3 table3.tex  (71Kb)

Savaglio S., Cristiani S., D’Odorico S., Fontana A., Giallongo E., Molaro P. Astron. Astrophys. 318, 347 (1997) The shape of the ionizing UV background at z 3.7 from the metal absorption systems of Q 0000-2619

J/A+A/318/416/  coravel elodie photom  (27Kb)

Pont F., Queloz D., Bratschi P., Mayor M. Astron. Astrophys. 318, 416 (1997) Rotation of the outer disc from classical Cepheids.

J/A+A/318/680/  angsiz.aux angsiz.tex  (37Kb)

Kayser R., Helbig P., Schramm T. Astron. Astrophys. 318, 680 (1997) A general and practical method for calculating cosmological distances

J/A+A/318/797/  table2  (9Kb)

Ripepi V., Barone F., Milano L., Russo G. Astron. Astrophys. 318, 797 (1997) Cepheid radii and the CORS method revisited

J/A+A/318/879/  appena.tex appenb.tex appenc.tex tablec1  (39Kb)

Men’shchikov A.B., Henning T. Astron. Astrophys. 318, 879 (1997) Radiation transfer in circumstellar disks

J/A+A/319/18/  appena table1  (31Kb)

Radecke H.-D. Astron. Astrophys. 319, 18 (1997) X-ray sources 1 degree from Seyfert galaxies.

J/A+A/319/33/  table3 table3.tex  (12Kb)

Arp H. Astron. Astrophys. 319, 33 (1997) Identification of X-ray sources 1^{∘}from Seyfert galaxies.

J/A+A/319/67/  table1 table2 tables.tex  (26Kb)

Huchtmeier W.K., Hopp U., Kuhn B. Astron. Astrophys. 319, 67 (1997) HI observations of dwarf galaxies in voids

J/A+A/319/184/  table1 table2  (23Kb)

Alcala J.M., Krautter J., Covino E., Neuhaeuser R., Schmitt J.H.M.M., Wichmann R. Astron. Astrophys. 319, 184 (1997) A study of the Chamaeleon star forming region from the ROSAT all-sky survey: II. The pre-main sequence population.

J/A+A/319/235/  table1 table2 tables.tex  (24Kb)

Bachiller R., Fuente A., Bujarrabal V., Colomer F., Loup C., Omont A., de Jong T. Astron. Astrophys. 319, 235 (1997) A survey of CN in circumstellar envelopes

J/A+A/319/318/ tablea1a tablea1b tablea1c tablea1d tablea2a tablea2b tablea2c tablea2d tablea2e tablea2f tablea2g tablea2h (36Kb) Pauzat F., Talbi D., Ellinger Y. Astron. Astrophys. 319, 318 (1997) The PAH hypothesis: a computational experiment on the combined effects of ionization and dehydrogenation on the IR signatures

J/A+A/319/401/  table1 table2 tables.tex  (24Kb)

Huchtmeier W.K. Astron. Astrophys. 319, 401 (1997) An HI survey of polar ring galaxies. II. The Effelsberg sample.

J/A+A/319/413/  table1  (78Kb)

Brinkmann W., Yuan W., and Siebert J. Astron. Astrophys. 319, 413 (1997) Broad-band energy distribution of ROSAT detected quasars. I. Radio-loud objects.

J/A+A/319/481/  table3.dat  (16Kb)

Mermilliod J.-C., Andersen J., Mayor M. Astron. Astrophys. 319, 481 (1997) Red giants in open clusters. VI. Orbits of ten spectroscopic binaries.

J/A+A/319/593/  table4 table5 table6 table7  (15Kb)

Molaro P., Bonifacio P., Castelli F., Pasquini L. Astron. Astrophys. 319, 593 (1997) New beryllium observations in low-metallicity stars

J/A+A/319/757/  table1 table2 table3  (16Kb)

Hoekstra H., Barthel P.D., Hes R. Astron. Astrophys. 319, 757 (1997) Beamed radio and far infrared emission in quasars and radio galaxies

J/A+A/319/850/  ter4iz ter4vi  (452Kb)

Ortolani S., Barbuy B., Bica E. Astron. Astrophys. 319, 850 (1997) High quality seeing V, I, and Gunn z imaging of Terzan 4: a blue horizontal branch bulge globular cluster

J/A+A/319/867/  table3a table3b  (39Kb)

Harmanec P., Hadrava P., Yang S., Holmgren D., North P., Koubsky P., Kubat J., Poretti E. Astron. Astrophys. 319, 867 (1997) Search for forced oscillations in binaries. I. The eclipsing and spectroscopic binary V436 Persei = 1 Persei

J/A+A/319/881/  table2  (18Kb)

Gomez A.E., Luri X., Grenier S., Prevot L., Mennessier M.O., Figueras F., Torra J. Astron. Astrophys. 319, 881 (1997) Absolute magnitudes and kinematics of barium stars

J/A+A/320/74/  table2  (14Kb)

Mermilliod J.-C., Bratschi P., Mayor P. Astron. Astrophys. 320, 74 (1997) Investigation of the Pleiades cluster. III. Additional corona members

J/A+A/320/79/  refs.dat table2.dat table4.dat tables.tex  (39Kb)

Malkov O.Yu., Piskunov A.E., Shpil’kina D.A. Astron. Astrophys. 320, 79 (1997) Mass-luminosity relation of low-mass stars

J/A+A/320/185/  table6.dat table7.dat  (19Kb)

Wichmann R., Krautter J., Covino E., Alcala J.M., Neuhaeuser R., Schmitt J.H.M.M. Astron. Astrophys. 320, 185 (1997) The T Tauri star population in the Lupus star forming region

J/A+A/320/405/  table2  (10Kb)

Sergeev S.G., Pronik V.I., Malkov Yu.F., Chuvaev K.K. Astron. Astrophys. 320, 405 (1997) [O III] λ5007 wing variations and changes of Fe II optical multiplets in NGC 5548

J/A+A/320/631/  table6-7 tables.tex  (22Kb)

Jopek T.J., Froeschle C. Astron. Astrophys. 320, 631 (1997) A stream search among 502TV meteor orbits. An objective approach

J/A+A/320/652/ fig5.ps fig8a.ps fig8b.ps fig9.ps fig11.ps fig12a.ps fig12b.ps table1 table1.ps table5.ps table6.ps table7 table7.ps (516Kb) Viateau B., Rapaport M. Astron. Astrophys. 320, 652 (1997) The Bordeaux meridian observations of asteroids. First determination of the mass of (11) Parthenope

J/A+A/320/757/  tables  (1.1 Mb)

Ferraro F.R., Carretta E., Corsi C.E., Fusi Pecci F., Cacciari C., Buonanno R., Paltrinieri B., Hamilton D. Astron. Astrophys. 320, 757 (1997) The stellar population of the globular cluster M 3. II. CCD photometry of additional 10,000 stars

J/A+A/320/776/  table.dat  (8Kb)

Lehmann I., Scholz R.-D. Astron. Astrophys. 320, 776 (1997) Tidal radii of the globular clusters M 5, M 12, M 13, M 15, M 53, NGC 5053 and NGC 5466 from automated star counts.

J/A+A/321/L33/  table1 table1.tex  (29Kb)

Jetsu L. Astron. Astrophys. 321, L33 (1997) The ”human” statistics of terrestrial impact cratering rate

J/A+A/321/L9/  table1 table2 tables.tex  (11Kb)

Bergeat J., Knapik A. Astron. Astrophys. 321, L9 (1997) The Barium stars in the Hertzsprung-Russell diagram

J/A+A/321/29/ fig1.ps fig2.ps fig3.ps fig4.ps fig5.ps fig6.ps fig7.ps fig8.ps fig9.ps fig10.ps fig11.ps fig12.ps fig13.ps table (945Kb) Olofsson K. Astron. Astrophys. 321, 29 (1997) Effects of star cluster evolution on emission line ratios and the application to H II galaxies. Including a calibration of the R_{23} vs. 12+log(O/H) relationship

J/A+A/321/123/  table1  (11Kb)

Cristiani S., Trentini S., La Franca F., Andreani P. Astron. Astrophys. 321, 123 (1997) The optical variability of QSOs. II. The wavelength dependence

J/A+A/321/236/  table3 table3.tex  (16Kb)

Knapik A., Bergeat J.

Astron. Astrophys. 321, 236 (1997) Interstellar extinction and the intrinsic spectral distribution of variable carbon stars

J/A+A/321/293/  table1a table1b table2  (19Kb)

Schilke P., Walmsley C.M., Pineau des Forets G., Flower D.R. Astron. Astrophys. 321, 293 (1997) SiO production in interstellar shocks.

J/A+A/321/363/  table3 table4 tables.tex  (21Kb)

Martin P., Belley J. Astron. Astrophys. 321, 363 (1997) Nebular gas abundances and mixing processes in the ringed galaxy NGC 4736

J/A+A/321/531/  optical.dat tables.tex uv.dat  (36Kb)

Taresch G., Kudritzki R.P., Hurwitz M., Bowyer S., Pauldrach A.W.A., Puls J., Butler K., Lennon D.J., Haser S.M. Astron. Astrophys. 321, 531 (1997) Quantitative analysis of the FUV, UV and optical spectrum of the O3 star HD 93129A

J/A+A/322/159/  table2  (7Kb)

Szymczak M., Engels D. Astron. Astrophys. 322, 159 (1997) H_{2}O maser emission from irregular variables

J/A+A/322/460/  table1  (14Kb)

Nordstrom B., Andersen J., Andersen M.I. Astron. Astrophys. 322, 460 (1997) Critical tests of stellar evolution in open clusters. II. Membership, duplicity, and stellar and dynamical evolution in NGC 3680

J/A+A/322/817/  table1  (11Kb)

Feuchtinger M.U., Dorfi E.A. Astron. Astrophys. 322, 817 (1997) Properties of theoretical RRab light curves

J/A+A/323/L49 See I/239/

J/A+A/323/71/  table1 table2p1 table2p2 table3 tables.tex  (56Kb)

Contini M. Astron. Astrophys. 323, 71 (1997) Line spectra of Liners. The contribution of the different ionization mechanisms.

J/A+A/323/113/  table5 table6  (31Kb)

Munari U., Rejkuba M., Hazen M., Mattei J., Schweitzer E., Luthardt R., Shugarov S., Yudin B.F., Popova A.A., Chugainov P. V., Sostero G., Lepardo A. Astron. Astrophys. 323, 113 (1997) The symbiotic star YY Her. I. Photometric history over 1890-1996.

J/A+A/323/139/  table3.dat  (7Kb)

Bouvier J., Rigaut F., Nadeau D. Astron. Astrophys. 323, 139 (1997) Pleiades low-mass binaries: do companions affect the evolution of protoplanetary disks?

J/A+A/323/337/  table1 table1.tex  (26Kb)

Andreon S., Davoust E., Heim T. Astron. Astrophys. 323, 337 (1997) The morphological segregation of galaxies in clusters. III. The distant cluster Cl 0939+4713 (Abell 851)

J/A+A/323/469/  table5 tablea1 tablea2 tablea3 tablea4 tablea5 tables.tex  (71Kb)

Bakker E.J., van Dishoeck E.F., Waters L.B.F.M., Schoenmaker T. Astron. Astrophys. 323, 469 (1997) Circumstellar C_{2}, CN, and CH^{+} in the optical spectra of post-AGB stars

J/A+A/323/707/  tab1-17 table18 tables.tex  (75Kb)

Erkens U., Appenzeller I., Wagner S. Astron. Astrophys. 323, 707 (1997) The nature of the FHIL winds from AGN

J/A+A/323/739/  table1.dat table2.dat table3.dat table4.dat tables.tex  (359Kb)

Brinkmann W., Siebert J., Feigelson E.D., Kollgaard R.I., Laurent-Muehleisen S.A., Reich W., Fuerst E., Reich P., Voges W., Truemper J., McMahon R. Astron. Astrophys. 323, 739 (1997)

Radio-loud AGN in the northern ROSAT All-Sky Survey. II. Multi-frequency properties of unidentified sources

J/A+A/323/775/  table2  (14Kb)

Mishurov Yu.N., Zenina I.A., Dambis A.K., Mel’nik A.M., Rastorguev A.S. Astron. Astrophys. 323, 775 (1997) Is the Sun located near the corotation circle?

J/A+A/323/881/  table1  (8Kb)

Mathys G., Lanz T. Astron. Astrophys. 323, 881 (1997) The variations of the Bp star HD 137509

J/A+A/324/91/  table4  (13Kb)

Mermilliod J.-C., Claria J.J., Andersen J., Mayor M. Astron. Astrophys. 324, 91 (1997) Red giants in open clusters: VII. Melotte 71

J/A+A/324/137/  table1 table2 table3  (54Kb)

North P., Studer M., Kunzli M. Astron. Astrophys. 324, 137 (1997) Eclipsing binaries with candidate CP stars. I. Parameters of the systems HD 143654, HD 184035 and HD 185257

J/A+A/324/366/  table3 table4 table5 table6 table7 table8 tables.tex  (128Kb)

Duriez L., Vienne A. Astron. Astrophys. 324, 366 (1997) Theory of motion and ephemerides of Hyperion

J/A+A/324/435/  appen  (26Kb)

Hill V. Astron. Astrophys. 324, 435 (1997) Chemical composition of six K supergiants in the Small Magellanic Cloud.

J/A+A/324/505/  table67  (11Kb)

Dubath P., Meylan G., Mayor M. Astron. Astrophys. 324, 505 (1997) Core velocity dispersions for 25 Galactic and 10 old Magellanic globular clusters

J/A+A/324/549/  note table3 table5  (23Kb)

Trullols E., Jordi C. Astron. Astrophys. 324, 549 (1997) Deep UVBRI photometry in IC 348

J/A+A/324/904/  figure3.ps table1 table2 table3 table4  (106Kb)

Malkov Yu.F., Pronik V.I., Sergeev S.G. Astron. Astrophys. 324, 904 (1997) Complex variability pattern in NGC 4151. I. Sequences on the line-continuum diagram

J/A+A/324/915/  table2 table3 table4  (38Kb)

Ferraro F.R., Paltrinieri B., Fusi Pecci F., Cacciari C., Dorman B., Rood R.T., Buonanno R., Corsi C.E., Burgarella D., Laget M. Astron. Astrophys. 324, 915 (1997) HST observations of Blue Stragglers stars in the core of the globular cluster M3

J/A+A/324/949/  table1 table1.ps table1.tex  (42Kb)

Zboril M., North P., Glagolevskij Yu.V., Betrix F. Astron. Astrophys. 324, 949 (1997) Properties of He-rich stars I. Their evolutionary state and helium abundance

J/A+A/324/965/  table4  (61Kb)

Hill G., Harmanec P., Pavlovski K., Bozic H., Hadrava P., Koubsky P., Ziznovsky J. Astron. Astrophys. 324, 965 (1997) Properties and nature of Be stars. 17. V360 Lac = HD 216200 is a B3e + F9IV: binary

J/A+A/324/1059/  table2 table3 table4 tables.tex  (39Kb)

Le Bertre T. Astron. Astrophys. 324, 1059 (1997) Optical and infrared observations of 23 carbon-rich stars. Modelling of the circumstellar dust shells

J/A+A/325/159/  table1 table2 table3 tables.tex  (31Kb)

Leinert C., Henry T., Glindemann A., McCArthy D.W. Jr. Astron. Astrophys. 325, 159 (1997)

A search for companions to nearby southern M dwarfs with near-infrared speckle interferometry

J/A+A/325/167/  table3.dat table4.dat table5.dat  (32Kb)

Raboud D., Cramer N., Bernasconi P.A. Astron. Astrophys. 325, 167 (1997) Geneva photometry in the young open cluster NGC 6231

J/A+A/325/255/  table1 table2 table3 table4 table6 table7 tables.tex  (116Kb)

Baudry A., Desmurs J.-F., Wilson T.L., Cohen R.J. Astron. Astrophys. 325, 255 (1997) A survey of star-forming regions in the 5 cm lines of OH

J/A+A/325/473/  table1 table2 tables.tex  (34Kb)

Huchtmeier W.K. Astron. Astrophys. 325, 473 (1997) HI-deficiency in Hickson compact groups of galaxies

J/A+A/325/647/  table1 table2  (19Kb)

Neuhauser R., Torres G., Sterzik M.F., Randich S. Astron. Astrophys. 325, 647 (1997) Optical high-resolution spectroscopy of ROSAT detected late-type stars south of the Taurus molecular clouds

J/A+A/325/693/  table4  (9Kb)

Malyuto V., Schmidt-Kaler T. Astron. Astrophys. 325, 693 (1997) Quantitative spectral classification based on photoelectric spectrum scanner measurements of F-K stars

J/A+A/325/987/  bmag table4.dat umag vmag  (200Kb)

Krzesinski J., Pigulski A. Astron. Astrophys. 325, 987 (1997) B-type pulsators in the open cluster NGC 884 (χPersei)

J/A+A/325/1115/  table1 table2  (44Kb)

Piters A.J.M., Schrijver C.J., Schmitt J.H.M.M., Rosso C., Baliunas S.L., Van Paradijs J., Zwaan C. Astron. Astrophys. 325, 1115 (1997) Coronal and chromospheric emission from cool stars in near-simultaneous ROSAT All-Sky Survey and Mount Wilson data

J/A+A/325/1163/  fig1.ps fig2.ps fig3.ps fig5.ps fig6.ps table5 table6  (535Kb)

Millar T.J., Macdonald G.H., Gibb A.G. Astron. Astrophys. 325, 1163 (1997) A 330-360 GHz spectral survey of G34.3+0.15. II. Chemical modelling.

J/A+A/326/211/  table1.dat  (12Kb)

Wichmann R., Sterzik M., Krautter J., Metanomski A., Voges W. Astron. Astrophys., 326, 211-217 (1997) T Tauri stars and the Gould Belt near Lupus.

J/A+A/326/249/  table4 table5 tables.tex  (19Kb)

Mauas P., Falchi A., Pasquini L., Pallavicini R. Astron. Astrophys. 326, 249 (1997) Chromospheric models of dwarf M stars

J/A+A/326/614/  table2  (27Kb)

Ortolani S., Bica E., Barbuy B. Astron. Astrophys. 326, 614 (1997) V, I photometry of the metal-rich bulge globular cluster Terzan 2

J/A+A/326/722/  table4  (18Kb)

Mennessier M.O., Luri X., Figueras F., Gomez A.E., Grenier S., Torra J., North P. Astron. Astrophys. 326, 722 (1997) Ba stars, galactic populations and evolution

J/A+A/326/751/  tablea1.dat tablea2.dat tablea3.dat tablea4.dat tablea5.dat  (64Kb)

Nissen P.E., Schuster W.J. Astron. Astrophys. 326, 751 (1997) Chemical composition of halo and disk stars with overlapping metallicities

J/A+A/326/780/  table1 table2 table3 tables.tex  (47Kb)

Barrado y Navascues D., Fernandez-Figueroa M.J., Garcia Lopez R.J., De Castro E., Cornide M. Astron. Astrophys. 326, 780 (1997) The age-mass relation for chromospherically active binaries. II. Lithium depletion in dwarf components

J/A+A/326/1044/  table5  (14Kb)

Jordi C., Ribas I., Torra J., Gimenez A. Astron. Astrophys. 326, 1044 (1997) Photometric versus empirical surface gravities of eclipsing binaries

J/A+A/326/1069/  table5 table6 table7 table9 table10 tables.tex  (102Kb)

Russell S.C. Astron. Astrophys. 326, 1069 (1997) More Light on RV Tau Variables

J/A+A/327/61/  table2.dat table2.tex  (15Kb)

Ghisellini G., Villata M., Raiteri C.M., Bosio, S., De Francesco G., Latini G., Maesano M., Massaro E., Montagni F., Nesci R., Tosti G., Fiorucci M., Pian E., Maraschi L., Treves A., Comastri A. and Mignoli M. Astron. Astrophys. 327, 61 (1997) Otical-IUE observations of the gamma-loud BL Lacertae object S5 0716+714: data and interpretation

J/A+A/327/145/  table1 table2  (30Kb)

Rostopchina A.N., Grinin V.P., Okazaki A., The P.S., Kikuchi S., Shakhovskoy D.N., Minikhulov N.Kh. Astron. Astrophys. 327, 145 (1997) Dust around young stars. Photopolarimetric activity of the classical Herbig Ae/Be star RR Tauri.

J/A+A/327/191/  table1 table3 table4 tables.tex  (43Kb)

Mikolajewska J., Acker A., Stenholm B. Astron. Astrophys. 327, 191 (1997) Spectrophotometric study of southern symbiotic stars

J/A+A/327/325/  table2.dat  (11Kb)

May J., Alvarez H., Bronfman L. Astron. Astrophys. 327, 325 (1997) Physical properties of molecular clouds in the southern outer Galaxy.

J/A+A/327/522/  fig/* table1.dat table2.dat table3.dat  (1.9 Mb)

Boselli A., Gavazzi G., Lequeux J., Buat V., Casoli F., Dickey J., Donas J. Astron. Astrophys. 327, 522 (1997) The molecular gas content of spiral galaxies in the Coma/A1367 supercluster

J/A+A/327/587/  table4  (43Kb)

Tomkin J., Edvardsson B., Lambert D.L., Gustafsson B. Astron. Astrophys. 327, 587 (1997) The rise and fall of the NaMgAl stars

J/A+A/327/736/  table1  (10Kb)

Stasinska G., Gorny S.K., Tylenda R. Astron. Astrophys. 327, 736 (1997) On the mass distribution of planetary nebulae central stars

J/A+A/327/755/  table1 table2  (10Kb)

Palla F., Codella C., Valdettaro R., Baffa C. Astron. Astrophys. 327, 755 (1997) Search for H_{2}O maser emission from high-latitude IRAS sources

J/A+A/327/952/  table5.ps table5.tex table5a.dat table5b.dat table5c.dat  (93Kb)

Stein P., Jerjen H., Federspiel M. Astron. Astrophys. 327, 952 (1997) Velocity structure of the dwarf galaxy population in the Centaurus cluster

J/A+A/327/1017/ (3Kb) Rovithis-Livaniou H., Rovithis P., Oprescu G., Dumitrescu A., Suran M.D. Astron. Astrophys. 327, 1017 (1997) The GO Cygni system: a near contact eclipsing binary

J/A+A/327/1039/  tab2-7  (29Kb)

Chabrier G., Baraffe I. Astron. Astrophys. 327, 1039 (1997) Structure and evolution of low-mass stars

J/A+A/327/1114/  table2 table2.tex  (60Kb)

Leto G., Pagano I., Buemi C.S., Rodono M. Astron. Astrophys. 327, 1114 (1997) Long-term photometric flares patrol (1967-1977) on EV Lac: a clear evidence of a longitude concentrated flaring activity in 1970

J/A+A/327/1177/  table1 table1.tex  (54Kb)

Wilson T.L., Filges L., Codella C., Reich W., Reich P. Astron. Astrophys. 327, 1177 (1997) Kinematics and electron temperatures in the core of Orion A

J/A+A/328/83/  tables  (14Kb)

Chabrier G., Mera D. Astron. Astrophys. 328, 83 (1997) Determination of the globular cluster and halo stellar mass functions and stellar and brown dwarf densities

J/A+A/328/167/  fig3.ps table1.dat  (2.8 Mb)

Braun J.M., Bomans D.J., Will J.-M., De Boer K.S. Astron. Astrophys. 328, 167 (1997) No stellar age gradient inside supergiant shell LMC4.

J/A+A/328/175/  tab10ab tab10cd tab10e  (161Kb)

Aoki W., Tsuji T. Astron. Astrophys. 328, 175 (1997) High resolution infrared spectroscopy of CN and NH lines: nitrogen abundance in oxygen-rich giants through K to late M

J/A+A/328/235/  table5 table5.tex  (12Kb)

Rodriguez E., Gonzalez-Bedolla S.F., Rolland A., Costa V., Lopez-Gonzalez M.J., Lopez de Coca P. Astron. Astrophys. 328, 235 (1997) Multiperiodicity in the atypical Delta Sct star AN Lyn

J/A+A/328/261/  table4  (19Kb)

Neuforge-Verheecke C., Magain P. Astron. Astrophys. 328, 261 (1997) Spectroscopic analysis of the Alpha Centauri System

J/A+A/328/349/  cm m05.dat cm m10.dat cm p00.dat cm p05.dat cm p10.dat table5  (66Kb)

Smalley B., Kupka F. Astron. Astrophys. 328, 349 (1997) The role of convection on the uvby colours of A, F, and G stars

J/A+A/328/551/  table1 table2  (19Kb)

Koubsky P., Harmanec P., Kubat J., Hubert A.-M., Bozic H., Floquet M., Hadrava P., Hill G., Percy J.R. Astron. Astrophys. 328, 551 (1997) Properties and nature of Be stars. XVIII. Spectral, light and colour variations of 4 Herculis

J/A+A/328/617/  table4.tex table4a table4b table4c  (25Kb)

Marechal P., Viala Y.P., Pagani L. Astron. Astrophys. 328, 617 (1997) Chemistry and rotational excitation of O2 in interstellar clouds. II - The ^{16}O^{18}O isotopomer

J/A+A/329/101/  table3  (33Kb)

Raboud D., Mermilliod J.-C. Astron. Astrophys. 329, 101 (1998) Investigation of the Pleiades cluster. IV. The radial structure.

J/A+A/329/219/  tab5-16.dat tab5-16.ps  (169Kb)

Pardo J.R., Cernicharo J., Gonzalez-Alfonso E., Bujarrabal V. Astron. Astrophys. 329, 219 (1998) Simultaneous observations of maser lines of ^{28}SiO in evolved stars

J/A+A/329/329/ angdyn97.ome angdyn97.psi example.f figdyn97.chi figdyn97.eps figdyn97.ome figdyn97.p figdyn97.phi figdyn97.psi figdyn97.sid figkin97.chi figkin97.eps figkin97.ome figkin97.p figkin97.phi figkin97.psi figkin97.sid rotdyn97.ome rotdyn97.psi smart97.chk smart97.doc smart97.f (8.6 Mb) Bretagnon P., Francou G., Rocher P., Simon J.L. Astron. Astrophys. 329, 329 (1998) SMART97: a new solution for the rotation of the rigid Earth

J/A+A/329/482/  table4.dat  (9Kb)

Hasinger G., Burg R., Giacconi R., Schmidt M., Truemper J., Zamorani G. Astron. Astrophys., 329, 482-494 (1998) The ROSAT Deep Survey. I. X-ray sources in the Lockman Field.

J/A+A/329/579/  table2aa.dat table2ab.dat table2b.dat  (43Kb)

Fabregat J., Adelman S.J. Astron. Astrophys. 329, 579 (1998)

uvby photometry of theta Coronae Borealis during 1994 and 1995

J/A+A/329/721/  table1 table2  (51Kb)

Sheminova V. A. Astron. Astrophys. 329, 721 (1998) Sensitivity indicators of Fraunhofer lines

J/A+A/329/943/  table3 table4 table5 table6 table7 table8  (32Kb)

Ng Y.K., Bertelli G. Astron. Astrophys. 329, 943 (1998) Revised ages for stars in the solar neighbourhood

J/A+A/330/108/  table1  (119Kb)

Yuan W., Brinkmann W., Siebert J., Voges W. Astron. Astrophys. 330, 108 (1998) Broad band energy distribution of ROSAT detected quasars. II. Radio-quiet objects

J/A+A/330/423/  tables.ps tables.tex  (35Kb)

Rampazzo R., Covino S., Trinchieri G., Reduzzi L. Astron. Astrophys. 330, 423 (1998) Testing the physical reality of binaries and compact groups. Properties of early-type galaxies in groups with diffuse X-ray emission

J/A+A/330/453/  m54.dat sagit1.dat sagit2.dat  (782Kb)

Marconi G., Buonanno R., Castellani M., Iannicola G., Molaro P., Pasquini L., Pulone L. Astron. Astrophys. 330, 453 (1998) The stellar content of the Sagittarius Dwarf Galaxy.

J/A+A/330/505/  table1.dat table2.dat tables.tex  (22Kb)

Blanco A., Borghesi A., Fonti S., Orofino V. Astron. Astrophys. 330, 505 (1998) Circumstellar emission from dust envelopes around carbon stars showing the silicon carbide feature

J/A+A/330/515/  table1 table1.tex  (63Kb)

Fernley J., Barnes T.G., Skillen I., Hawley S.L., Hanley C.J., Evans D.W., Solano E., Garrido R. Astron. Astrophys. 330, 515 (1998) The absolute magnitudes of RR Lyrae from HIPPARCOS parallaxes and proper motions

J/A+A/330/676/  table5 table5.tex  (95Kb)

Willacy K., Cherchneff I. Astron. Astrophys. 330, 676 (1998) Silicon and sulfur chemistry in the inner wind of IRC+10216

J/A+A/330/881/  table3.tex table3a table3b  (15Kb)

Amram P., Mendes de Oliveira C., Boulesteix J., Balkowski C. Astron. Astrophys. 330, 881 (1998) The Hαkinematics of the Cartwheel galaxy

J/A+A/330/990/  table2 table2.tex  (36Kb)

Giovannetti P., Caux E., Nadeau D., Monin J.-L. Astron. Astrophys. 330, 990 (1998) Deep optical and near infrared imaging photometry of the Serpens cloud core

J/A+A/331/70/  appen.dat  (63Kb)

Guarnieri M.D., Ortolani S., Montegriffo P., Renzini B., Barbuy B., Bica E., Moneti A. Astron. Astrophys. 331, 70 (1998) Infrared Array Photometry of Bulge Globular Clusters. I. Combined ground based JK and HST VI Photometry of NGC 6553

J/A+A/331/81/  table2.dat  (41Kb)

Perryman M.A.C., Brown A.G.A., Lebreton Y., Gomez A., Turon C., Cayrel de Strobel G., Mermilliod J.C., Robichon N., Kovalevsky J., Crifo F. Astron. Astrophys. 331, 81 (1998) The Hyades: distance, structure, dynamics, and age

J/A+A/331/581/  table2.dat table4.dat  (26Kb)

Delfosse X., Forveille T., Perrier C., Mayor M. Astron. Astrophys. 331, 581 (1998) Rotation and chromospheric activity in field M dwarfs

J/A+A/331/633/  table1.dat table2.dat table3.dat tables.tex  (29Kb)

Solano E., Paunzen E.

Astron. Astrophys. 331, 633 (1998) Identification of Lambda Bootis stars using IUE spectra. I. Low Resolution Data

J/A+A/331/669 See VIII/66/

J/A+A/331/949/  fig/* table1.dat table1.ps table1.tex  (579Kb)

Moffat A.F.J., Marchenko S.V., Seggewiss W., van der Hucht K.A., Schrijver H., Stenholm B., Lundstrom I., Setia Gunawan D.Y.A., Sutantyo W., van den Heuvel E.P.J., De Cuyper J.-P., Gomez A.E. Astron. Astrophys. 331, 949 (1998) Wolf-Rayet stars and O-star runaways with HIPPARCOS. I. Kinematics. (Based on data from the ESA Hipparcos astrometry satellite)

J/A+A/332/L53/  table1.dat table1.ps table1.tex  (32Kb)

Bergeat J., Knapik A., Rutily B. Astron. Astrophys. 332, L53 (1998) The PL relation of galactic carbon LPVs. The distance modulus to LMC

J/A+A/332/410/  table23.dat  (14Kb)

Hutsemekers D. Astron. Astrophys. 332, 410 (1998) Evidence for very large-scale coherent orientations of quasar polarization vectors

J/A+A/332/575/  catalog.dat  (58Kb)

Belikov A.N., Hirte S., Meusinger H., Piskunov A.E., Schilbach E. Astron. Astrophys. 332, 575 (1998) The fine structure of the Pleiades luminosity function and pre-main sequence evolution

J/A+A/332/586/  table1.dat table1.tex  (220Kb)

Fritz T., Bruch A. Astron. Astrophys. 332, 586 (1998) Studies of the flickering in cataclysmic variables. IV. Wavelet transforms of flickering light curves

J/A+A/332/721/  abund.dat den.dat fluxes.dat table1.tex table2.tex tnii.dat toiii.dat  (213Kb)

Perinotto M., Corradi R. Astron. Astrophys. 332, 721 (1998) The chemical structure of bipolar planetary nebulae. II. 13 objects

J/A+A/333/L35/  table1.dat table2.dat table3.dat tables.ps tables.tex  (40Kb)

Antonello E. Astron. Astrophys. 333, L35 (1998) The possible effects of an unusual resonance in very long period Cepheids

J/A+A/333/205/  table.dat  (12Kb)

Frasca A., Marilli E., Catalano S. Astron. Astrophys. 333, 205 (1998) Photospheric and chromospheric activity in the late-type giant component of the evolved binary system HD 185510

J/A+A/333/231/  table1.dat table2.dat table3.dat table4.dat table5.dat table6.dat  (94Kb)

Bessell M.S., Castelli F., Plez B. Astron. Astrophys. 333, 231 (1998) Model atmospheres broad-band colors, bolometric corrections and temperature calibrations for O - M stars

J/A+A/333/471/  crossid.dat table5.dat  (369Kb)

Galadi-Enriquez D., Jordi C., Trullols E., Ribas I. Astron. Astrophys. 333, 471 (1998) The overlapping open clusters NGC 1750 and NGC 1758. I. UBVRI-CCD photometry

J/A+A/333/479/  appen.ps appen.tex cback.ps field1ms.ps xfig.ps  (323Kb)

Saviane I., Piotto G., Fagotto F., Zaggia S., Capaccioli M., Aparicio A. Astron. Astrophys. 333, 479 (1998) The Galactic globular cluster NGC 1851: its dynamical and evolutionary properties

J/A+A/333/619/  tablea1.dat tablea2.dat  (22Kb)

Preibisch T., Guenther E., Zinnecker H., Sterzik M., Frink S., Roeser S. Astron. Astrophys. 333, 619 (1998) A lithium-survey for pre-main sequence stars in the upper Scorpius OB association

J/A+A/333/658/  table7.dat table7.tex  (49Kb)

Pereira C.B., Landaberry S.J.C., Junqueira S. Astron. Astrophys. 333, 658 (1998) Spectroscopic observations of D-type symbiotic stars

J/A+A/333/897/  table2.dat table3.dat  (42Kb)

Raboud D., Mermilliod J.-C. Astron. Astrophys. 333, 897 (1998) Evolution of mass segregation in open clusters: some observational evidences

J/A+A/333/952/  table1.dat  (14Kb)

Adelman S.J. Astron. Astrophys. 333, 952 (1998) BVRI photometry of the extrinsic S Star HR 1105

J/A+A/334/609/  table3.dat table4.dat tables.tex  (39Kb)

Guglielmo F., Le Bertre T., Epchtein N. Astron. Astrophys. 334, 609 (1998) Infrared carbon stars: new identifications and their space distribution in the Galaxy

J/A+A/334/863/  table2.dat  (8Kb)

Berdyugina S.V., Jankov S., Ilyin I., Tuominen I., Fekel F.C. Astron. Astrophys. 334, 863 (1998) The active RS Canum Venaticorum binary II Pegasi. I. Stellar and orbital parameters.

J/A+A/334/987/  appen.ps appen.tex appena.dat appenb.dat  (52Kb)

Venn K.A., Smartt S.J., Lennon D.J., Dufton P.L. Astron. Astrophys. 334, 987 (1998) New identifications for blue objects towards the Galactic center: post-AGB stars, Be/disk stars, and others

J/A+A/335/L65 See I/250/

J/A+A/335/85/  table1.dat table2.dat  (18Kb)

Oestlin G., Bergvall N., Roennback J. Astron. Astrophys. 335, 85 (1998) Globular clusters in the blue compact galaxy ESO 338-IG04 (Tololo 1924-416), as tracers of the star formation history. Results from HST/WFPC2 observations.

J/A+A/335/183/  table3.dat table4.dat table5.dat tables.ps tables.tex  (63Kb)

Queloz D., Allain S., Mermilliod J.-C., Bouvier J., Mayor M. Astron. Astrophys. 335, 183 (1998) The rotational velocity of low-mass stars in the Pleiades cluster

J/A+A/335/467/  notes.dat table2.dat table3.dat  (38Kb)

Thomas H.-C., Beuermann K., Reinsch K., Schwope A.D., Truemper J., Voges W. Astron. Astrophys. 335, 467 (1998) Identification of soft high galactic latitude RASS X-ray sources. I. A complete count-rate limited sample.

J/A+A/335/565/  table1.dat table2.dat  (13Kb)

Hubert A.M., Floquet M. Astron. Astrophys. 335, 565 (1998) Investigation of the variability of bright Be stars using Hipparcos photometry

J/A+A/335/647/  table1.dat table2.dat table3.dat  (0Kb)

Claret A. Astron. Astrophys. 335, 647 (1998) Very low mass stars: non-linearity of the limb-darkening laws

J/A+A/335/823/  table4.dat table6.dat tables.ps tables.tex  (142Kb)

Tantalo R., Chiosi C., Bressan A., Marigo P., Portinari L. Astron. Astrophys. 335, 823 (1998) Spectro-photometric evolution of elliptical galaxies. III. Infall models with gradients in mass density and star formation

J/A+A/335/912/  tables.dat tables.ps tables.tex  (66Kb)

Kelm B., Focardi P., Palumbo G.G.C Astron. Astrophys. 335, 912 (1998) Seyfert galaxies and environment: Seyferts in galaxy pairs and groups

J/A+A/335/973/  table2.dat  (11Kb)

Wade G.A., Hill G.M., Adelman S.J., Manset N., Bastien P. Astron. Astrophys. 335, 973 (1998) Magnetic field models for A and B stars. V. The magnetic field and photometric variations of 84 UMa

J/A+A/336/35/  table1.dat  (8Kb)

Jaaniste J., Tago E., Einasto M., Einasto J., Andernach H., Mueller V. Astron. Astrophys. 336, 35 (1998)

The supercluster-void network. IV. The shape and orientation of superclusters.

J/A+A/336/263/ fig12.ps fig13.ps fig14.ps fig15.ps fig16.ps fig17.ps table3.dat table4.dat tables.tex (940Kb) Bakker E.J., Lambert D.L., van Winckel H., McCarthy J.K., Waelkens C., Gonzalez G. Astron. Astrophys. 336, 263 (1998) Spectral variability of the binary HR 4049

J/A+A/336/339/ fig.dat map/* note1a.dat note1b.dat table1a.dat table1b.dat table2a.dat table2b.dat (816Kb) Molinari S., Brand J., Cesaroni R., Palla F., Palumbo G.G.C. Astron. Astrophys. 336, 339 (1998) A search for precursors of ultracompact HII regions in a sample of luminous IRAS sources. II. VLA observations.

J/A+A/336/667/  models.dat npgrid/* table1.dat table2.dat  (1.9 Mb)

Stasinska G., Richer M.G., Mc Call M.L. Astron. Astrophys. 336, 667 (1998) The planetary nebulae populations in five galaxies: abundance patterns and evolution

J/A+A/336/920/  table2.dat table3.dat table4.dat table5.dat  (49Kb)

Rodriguez E., Claret A., Garcia J.M., Zerbi F.M., Garrido R., Martin S., Akan C., Luedeke K., Keskin V., Ibanoglu C., Evren S., Tunca Z., Pekunlu R., Paparo M., Nuspl J., Krisciunas K., Jiang S.Y. Astron. Astrophys. 336, 920 (1998) HD 163151: a new W UMa type system

J/A+A/336/953/  table3.dat table4.dat  (53Kb)

Gomez A.E., Luri X., Grenier S., Figueras F., North P., Royer F., Torra J., Mennessier M.O. Astron. Astrophys. 336, 953 (1998) The HR diagram form Hipparcos data. Absolute magnitudes and kinematics of Bp-Ap stars.

J/A+A/337/17/  tables.dat  (14Kb)

Andreuzzi G., Bardelli S., Scaramella R., Zucca E. Astron. Astrophys. 337, 17 (1998) A redshift survey between the clusters of galaxies A 548 and A 3367.

J/A+A/337/125/  table1.dat table2.dat  (17Kb)

Galadi-Enriquez D., Jordi C., Trullols E. Astron. Astrophys. 337, 125 (1998) The overlapping open clusters NGC 1750 and NGC 1758. III. Cluster-field segregation and clusters physical parameters.

J/A+A/337/183/  table3.dat table5.dat tables.ps tables.tex  (49Kb)

Bolton C.T., Harmanec P., Lyons R.W., Odell A.P., Pyper D.M. Astron. Astrophys. 337, 183 (1998) HD 37017 = V1046 Ori. A double-lined spectroscopic binary with a B2e He-strong magnetic primary

J/A+A/337/403/  tab1-3.dat  (146Kb)

Baraffe I., Chabrier G., Allard F., Hauschildt P. Astron. Astrophys. 337, 403 (1998) Evolutionary models for solar metallicity low-mass stars: mass-magnitude relationships and color-magnitude diagrams

J/A+A/337/779/  table1.dat  (41Kb)

Kienzle F., Burki G., Burnet M., Meynet G. Astron. Astrophys. 337, 779 (1998) The pulsating yellow supergiant V810 Centauri

J/A+A/338/340/  table7.dat  (47Kb)

Mueller T.G., Lagerros J.S.V. Astron. Astrophys. 338, 340 (1998) Asteroids as far-infrared photometric standards for ISOPHOT

J/A+A/338/581/  table2.dat  (10Kb)

Schultheis M., Ng Y.K., Hron J., Kerschbaum F. Astron. Astrophys. 338, 581 (1998) Field #3 of the Palomar-Groningen survey. II. Near-infrared photometry of semiregular variables.

J/A+A/338/813/  table1.dat table1.ps table1.tex  (79Kb)

Wu X.-P., Fang L.-Z., Xu W. Astron. Astrophys. 338, 813 (1998) Updating the σ-T relationship for galaxy clusters

J/A+A/338/1006/  table5.dat  (21Kb)

Iijima T., Rosino L., Della Valle M. Astron. Astrophys. 338, 1006 (1998) Spectral evolution of Nova (V723) Cassiopeiae 1995: pre-maximum stage.

J/A+A/338/1066/ bri0021.dat grh2208.dat j0020-44.dat j0021-42.dat j0205-11.dat j1058-15.dat j1228-15.dat j2040-32.dat j2052-55.dat j2146-21.dat stars.dat vb10.dat (212Kb) Tinney C.G., Delfosse X., Forveille T., Allard F. Astron. Astrophys., 338, 1066-1072 (1998) Optical spectroscopy of DENIS Mini-Survey Brown Dwarf Candidates.

J/A+A/339/34/  table1.dat  (38Kb)

de Ruiter H.R., Parma P., Stirpe G.M., Perez-Fournon I., Gonzalez-Serrano I., Rengelink R.B., Bremer M.N. Astron. Astrophys. 339, 34 (1998) Bright galaxies from WENSS. I The minisurvey

J/A+A/339/123/ table5.dat table6.dat table7.dat table8.dat table9.dat table10.dat table11.dat (3.4 Mb) Driebe T., Schoenberner D., Bloeckera T., Herwig F. Astron. Astrophys. 339, 123 (1998) The evolution of helium white dwarfs. I. The companion of the millisecond pulsar PSR J1012+5307

J/A+A/339/165/  table2.dat  (51Kb)

Cowley C.R., Mathys G. Astron. Astrophys. 339, 165 (1998) Line identifications and preliminary abundances from the red spectrum of HD 101065 (Przybylski’s star)

J/A+A/339/423/  stars.dat table4a.dat table4b.dat  (24Kb)

Mermilliod J.-C., Mathieu R.D., Latham D.W., Mayor M. Astron. Astrophys. 339, 423 (1998) Red giants in open clusters: VIII. NGC 752

J/A+A/339/858/  table1.dat table2.dat  (44Kb)

Di Benedetto G. P. Astron. Astrophys. 339, 858 (1998) Towards a fundamental calibration of stellar parameters of A, F, G, K dwarfs and giants

J/A+A/340/196/  table1.dat table2.dat  (26Kb)

Rodriguez E., Claret A., Sedano J.L., Garcia J.M., Garrido R. Astron. Astrophys. 340, 196 (1998) The eclipsing binary system AB Cas: binarity and pulsation

J/A+A/340/300/  table1.dat  (14Kb)

Raassen A.J.J., Uylings P.H.M. Astron. Astrophys. 340, 300 (1998) On de determination of the solar iron abundance using Fe II lines

J/A+A/340/371/  table2.dat table3.dat  (12Kb)

Hutsemekers D., Lamy H., Remy M. Astron. Astrophys. 340, 371 (1998) Polarization properties of a sample of broad absorption line and gravitationally lensed quasars

J/A+A/341/L79/  table1.dat table1.ps table1.tex  (61Kb)

Terranegra L., Morale F., Spagna A., Massone G., Lattanzi M.G. Astron. Astrophys. 341, L79 (1999) Proper motions of faint ROSAT WTT stars in the Chamaeleon region.

J/A+A/341/98/  table2.dat  (116Kb)

Selman F., Melnick J., Bosch G., Terlevich R. Astron. Astrophys. 341, 98 (1999) The ionizing cluster of 30 Doradus. I. Internal reddening from NTT photometry.

J/A+A/341/121/  notes.dat table1.dat table3.dat table4.dat table6.dat  (62Kb)

Soederhjelm S. Astron. Astrophys. 341, 121 (1999) Visual binary orbits and masses post Hipparcos

J/A+A/341/163/  notes.dat table3.dat table3.ps table3.tex  (147Kb)

Hearty T., Magnani L., Caillault J.-P., Neuhaeuser R., Schmitt J.H.M.M., Stauffer J. Astron. Astrophys. 341, 163 (1999) A search for star formation in the translucent clouds MBM 7 and MBM 55

J/A+A/341/469/  table.dat  (9Kb)

Carrier F., Burki G., Richard C. Astron. Astrophys. 341, 469 (1999) Geneva photometry of the open cluster NGC 2451 and its exceptional Be star HR 2968

J/A+A/341/553/  table2.dat table3.dat  (31Kb)

Petrov P.P., Zajtseva G.V., Efimov Yu.S., Duemmler R., Ilyin I.V., Tuominen I., Shcherbakov V.A. Astron. Astrophys. 341, 553 (1999) Brightening of the T Tauri star RY Tau in 1996. Photometry, polarimetry and high resolution spectroscopy

J/A+A/341/641/  catalog.dat  (136Kb)

Arnouts S., D’Odorico S., Cristiani S., Zaggia S., Fontana A., Giallongo E. Astron. Astrophys. 341, 641 (1999) The NTT SUSI Deep Field.

J/A+A/341/709/  table5.dat table6.dat table7.dat table8.dat table9.dat tables.ps tables.tex  (85Kb)

Lindner U., Fritze - von Alvensleben U., Fricke K.J. Astron. Astrophys. 341, 709 (1999) Chemically consistent evolution of galaxies: Spiral galaxy models compared to DLA systems

J/A+A/341/751/  table2.dat table3.dat tables.ps tables.tex  (179Kb)

Micela G., Sciortino S., Harnden, F.R.Jr., Kashyap V., Rosner R., Prosser C.F., Damiani F., Stauffer J., Caillault J.-P. Astron. Astrophys. 341, 751 (1999) Deep ROSAT HRI observations of the Pleiades

J/A+A/341/818/  data.dat table1.dat  (34Kb)

Kienzle F., Moskalik P., Bersier D, Pont F. Astron. Astrophys. 341, 818 (1999) Structural Properties of s-Cepheid Velocity Curves. Constraining the location of the ω_{4}=2ω_{1} resonance.

J/A+A/342/1/  table1.dat table1.ps table1.tex  (69Kb)

Ramella M., Zamorani G., Zucca E., Stirpe G.M., Vettolani G., Balkowski C., Blanchard A., Cappi A., Cayatte V., Chincarini G., Collins C., Guzzo L., MacGillivray H., Maccagni D., Maurogordato S., Merighi R., Mignoli M., Pisani A., Proust D., Scaramella R. Astron. Astrophys. 342, 1 (1999) The ESO Slice Project (ESP) galaxy redshift survey: VI. Groups of Galaxies.

J/A+A/342/313/  3/7um.dat 3/15um.dat isocat.dat  (25Kb)

Aussel H., Cesarsky C.J., Elbaz D., Starck J.L. Astron. Astrophys. 342, 313 (1999) Provisional ISOCAM observations of the Hubble Deep Field reduced with the PRETI method

J/A+A/342/378/  refs.dat table1.dat table1.ps table1.tex  (172Kb)

Gurvits L.I., Kellermann K.I., Frey S. Astron. Astrophys. 342, 378 (1999) The ”angular size - redshift” relation for compact radio structures in quasars and radio galaxies

J/A+A/342/671/ gaskin.dat halpha.dat nii6548.dat nii6583.dat sii6716.dat sii6730.dat stelkin.dat (76Kb) Corsini E.M., Pizzella A., Sarzi M., Cinzano P., Vega Beltran J.C., Funes J.G.S.J., Bertola F., Persic M., Salucci P. Astron. Astrophys. 342, 671 (1999) Dark Matter in early-type spiral galaxies: the case of NGC 2179 and of NGC 2775

J/A+A/342/736/  table2.dat table2.ps table2.tex  (23Kb)

Stelzer B., Wilms J., Staubert R., Gruber D., Rothschild R. Astron. Astrophys. 342, 736 (1999) Evolution of spectral parameters during a pre-eclipse dip of Her X-1

J/A+A/342/773/  table3.dat table4.dat  (12Kb)

Bergeat J., Knapik A., Rutily B. Astron. Astrophys. 342, 773 (1999) Dust extinction and intrinsic SEDs of carbon-rich stars. II. The hot carbon stars

J/A+A/342/881/  table1.dat table2.dat tables.ps tables.tex  (59Kb)

Goriely S. Astron. Astrophys. 342, 881 (1999) Uncertainties in the solar system r-abundance distribution

J/A+A/343/86/  table1.dat  (23Kb)

Gavazzi G., Boselli A. Astron. Astrophys. 343, 86 (1999)

On the local radio luminosity function of galaxies. I: the Virgo cluster

J/A+A/343/93/  table1.dat  (54Kb)

Gavazzi G., Boselli A. Astron. Astrophys. 343, 93 (1999) On the local radio luminosity function of galaxies. II: Environmental dependences among late-type galaxies

J/A+A/343/446/  tablea1.dat  (10Kb)

Domingo A, Figueras F. Astron. Astrophys. 343, 446 (1999) Late A-type stars: new Stroemgren photometric calibrations of absolute magnitudes from Hipparcos

J/A+A/343/806/  table2.dat table3.dat tables.ps tables.tex  (51Kb)

Munari U., Tomasella L. Astron. Astrophys. 343, 806 (1999) Kinematics and binaries in young stellar aggregates. I. The trapezium system BD+00 1617 in Bochum 2

J/A+A/343/841/  table5.dat  (15Kb)

Corradi R.L.M., Brandi E., Ferrer O.E., Schwarz H.E. Astron. Astrophys. 343, 841 (1999) A narrowband imaging survey of symbiotic stars

J/A+A/343/847/  table2.dat  (11Kb)

Van Genderen A.M., Sterken C., De Groot M., Burki G. Astron. Astrophys. 343, 847 (1999) Photometric behaviour of ηCarinae, a celestial Chinese lantern: 1974-1998.

J/A+A/343/894/  table.dat  (19Kb)

Djurasevic G., Zakirov M., Erkapic S. Astron. Astrophys. 343, 894 (1999) A photometric study of the W UMa-type system RZ Tauri

J/A+A/344/263/  table5.dat  (11Kb)

Knapik A., Bergeat J., Rutily B. Astron. Astrophys. 344, 263 (1999) Dust extinction and intrinsic SEDs of carbon-rich stars. III. The Miras, CS and SC stars

J/A+A/344/521/  table1.dat table2.dat tables.ps tables.tex  (51Kb)

Haberl F., Pietsch W. Astron. Astrophys. 344, 521 (1999) Variable X-ray sources in the LMC region observed by the ROSAT PSPC.

J/A+A/344/943/  table1.dat  (45Kb)

Deharveng L., Zavagno A., Nadeau D., Caplan J., Petit M. Astron. Astrophys. 344, 943 (1999) Sh 138: a compact HII region excited by a very young cluster

J/A+A/345/419/  table2.dat  (7Kb)

Longhetti M., Bressan A., Chiosi C., Rampazzo R. Astron. Astrophys. 345, 419 (1999) Star formation history of early-type galaxies in low density environments. V. Blue line-strength indices for the nuclear region

J/A+A/345/430/  appendix.dat  (44Kb)

Hill V. Astron. Astrophys. 345, 430 (1999) Abundances of cool supergiants in the SMC young cluster NGC 330.

J/A+A/345/471/  table1.dat tablea1.dat tablea2.dat  (16Kb)

Robichon N., Arenou F., Mermilliod J.-C., Turon C. Astron. Astrophys. 345, 471 (1999) Open clusters with Hipparcos. I. Mean astrometric parameters.

J/A+A/345/505/  curves/* table4.dat  (103Kb)

Krzesinski J., Pigulski A., Kolaczkowski Z. Astron. Astrophys. 345, 505 (1999) B-type pulsators in the central region of NGC 869 (h Persei)

J/A+A/345/521/  table3.dat table5.dat  (47Kb)

Flaccomio E., Micela G., Sciortino S., Favata F., Corbally C., Tomaney A. Astron. Astrophys. 345, 521 (1999) BVRI photometry of the star-forming region NGC 2264: the initial mass function and star-forming rate

J/A+A/345/629/  table1.dat table1.ps table1.tex  (54Kb)

Maciel W.J., Quireza C. Astron. Astrophys. 345, 629 (1999) Abundance gradients in the outer galactic disk from planetary nebulae

J/A+A/346/586/  table1.dat table2.dat  (2.4 Mb)

Perez Hernandez F., Claret A., Hernandez M.M., Michel E. Astron. Astrophys. 346, 586 (1999) Photometric parameters for rotating models of A- and F-type stars

J/A+A/346/785/  table1.dat table2.dat tables.ps tables.tex  (73Kb)

Sfeir D., Lallement R., Crifo F., Welsh B.Y. Astron. Astrophys. 346, 785 (1999) Mapping the contours of the Local Bubble: Preliminary results

J/A+A/346/955/ 60-90p.dat 60-90u.dat 90-120p.dat 90-120u.dat 120-150p.dat 120-150u.dat 150u.dat (52Kb) Leroy J.L. Astron. Astrophys. 346, 955 (1999) Interstellar dust and magnetic field at the boundaries of the Local Bubble Analysis of polarimetric data in the light of Hipparcos parallaxes

J/A+A/347/69/  table4.dat table5.dat  (11Kb)

Rolleston W.R.J., Hambly N.C., Keenan F.P., Dufton P.L., Saffer R.A. Astron. Astrophys. 347, 69 (1999) Early-type stars in the Galactic halo from the Palomar-Green Survey. II: A sample of distant, apparently young Population I stars

J/A+A/347/164/  table3.dat  (13Kb)

Wade G., Mathys G., North P. Astron. Astrophys. 347, 164 (1999) The Ap spectroscopic binary HD 59435 revisited

J/A+A/347/348/  table2.dat  (6Kb)

Grevesse N., Sauval A.J. Astron. Astrophys. 347, 348 (1999) The solar abundance of iron and the photospheric model

J/A+A/347/583/  table3a.dat table3b.dat table3c.dat  (23Kb)

Georgiev L.N., Koenigsberger G., Ivanov M.M., St-Louis N., Cardona O. Astron. Astrophys. 347, 583 (1999) Long-term coherent variations in the WR system EZ CMa : The binary scenario revisited

J/A+A/347/841/  table1.dat table1.ps table1.tex  (35Kb)

Heydari-Malayeri M., Charmandaris V., Deharveng L., Rosa M.R., Zinnecker H. Astron. Astrophys. 347, 841 (1999) HST observations of the very young SMC ”blob” N88A

J/A+A/348/184/ (3Kb) Rovithis P., Rovithis-Livaniou H., Suran M.D., Fragoulopoulou E., Skopal A. Astron. Astrophys. 348, 184 (1999) The first photometric study of the binary star WZ Cygni.

J/A+A/348/897/  tablea1.dat tablea2.dat tablea3.dat tablea4.dat tablea5.dat tablea6.dat  (27Kb)

Lachaume R., Dominik C., Lanz T., Habing H.J. Astron. Astrophys. 348, 897 (1999) Age determinations of main-sequence stars: combining different methods.

J/A+A/349/L69/  catalog.dat  (49.7 Mb)

Schultheis M., Ganesh S., Simon G., Omont A., Alard C., Borsenberger J., Copet E., Epchtein N., Fouque P., Habing H. Astron. Astrophys. 349, L69 (1999) Interstellar extinction towards the inner Galactic Bulge.

J/A+A/349/236 See II/240/

J/A+A/349/389 See IX/10A/

J/A+A/349/448/  table4.dat table7.dat  (239Kb)

Sanner J., Geffert M., Brunzendorf J., Schmoll J. Astron. Astrophys. 349, 448 (1999) Photometric and kinematic studies of open star clusters. I. NGC 581 (M 103).

J/A+A/349/515/  table3.dat table3a.tex table3b.tex table3c.tex  (600Kb)

Borkovits T., Biro I.B. Astron. Astrophys. 349, 515 (1999) Photometric analysis of the neglected eclipsing binary system DL Cygni

J/A+A/349/619/  ubvri.dat  (29Kb)

Bouvier J., Chelli A., Allain S., Carrasco L., Costero R., Cruz-gonzalez I., Dougados C., Fernandez M., Martin E.L., Menard F., Mennessier C., Mujica R., Recillas E., Salas L., Schmidt G., Wichmann R. Astron. Astrophys., 349, 619-635 (1999) Magnetospheric accretion onto the T Tauri star AA Tauri. I. Constraints from multisite spectrophotometric monitoring.

J/A+A/350/513/  table5.dat table5.ps table5.tex  (51Kb)

Kahanpaa J., Jetsu L., Alha L., Ezhkova O.V., Grankin K.N., Hackman T., Ibragimov M.A., Kokko M., Kondratiev V.B., Laakso T., Lindroos O., Mel’nikov S.Yu., Nikula P.O., Schultz J., Shevchenko V.S., Torppa J., Tuominen I., Tyynela J., Virtanen J. Astron. Astrophys. 350, 513 (1999) Time series analysis of V368 Cephei photometry

J/A+A/350/566/  table3.dat  (17Kb)

Bozic H., Ruzdjak D., Sudar D. Astron. Astrophys. 350, 566 (1999) Long-term changes of the Be star OT Geminorum

J/A+A/350/805/  fits/* object.dat ps/*  (2.8 Mb)

Grupe D., Beuermann K., Mannheim K., Thomas H.-C. Astron. Astrophys. 350, 805 (1999) New bright soft X-ray selected ROSAT AGN. II. Optical emission line properties

J/A+A/350/883/  list.dat table1.dat table2.dat table12.ps table12.tex  (37Kb)

Olofsson G., Huldtgren M., Kaas A.A., Bontemps S., Nordh L., Abergel A., Andre P., Boulanger F., Burgdorf M., Casali M.M., Cesarsky C.J., Davies J., Falgarone E., Montmerle T., Perault M., Persi P., Prusti T., Puget J.L., Sibille F. Astron. Astrophys. 350, 883 (1999) ISOCAM observations of the R CrA star formation region.

J/A+A/350/955/  abundcor.dat bound.dat levels.dat  (78Kb)

Gratton R.G., Carretta E., Eriksson K., Gustafsson B. Astron. Astrophys. 350, 955 (1999) Abundances of light elements in metal-poor stars. II. Non-LTE abundance corrections

J/A+A/351/212/  table1.dat table3.dat tables.ps tables.tex  (101Kb)

Jetsu L., Pelt J., Tuominen I. Astron. Astrophys. 351, 212 (1999) Time series analysis of V1794 Cygni long-term photometry

J/A+A/351/526/  table2.dat  (70Kb)

Ripepi V., Brocato E., Castellani V. Astron. Astrophys. 351, 526 (1999) CCD photometry of young MC clusters. I: NGC 2172.

J/A+A/352/39/  table1.dat table2.dat table3.dat table4.dat  (68Kb)

Woudt P.A., Kraan-Korteweg R.C., Fairall A.P. Astron. Astrophys. 352, 39 (1999) Extragalactic Large-Scale Structures behind the Southern Milky Way. III. Redshifts Obtained at the SAAO in the Great Attractor region

J/A+A/352/117/  table3.dat table4.dat  (26Kb)

Romano D., Matteucci F., Molaro P., Bonifacio P. Astron. Astrophys. 352, 117 (1999) The Galactic lithium evolution revisited

J/A+A/352/228/  north/* south/* table1.dat table1.ps table1.tex  (61.3 Mb)

Magnier E.A., Volp A.W., Laan K., Van den Ancker M.E., Waters L.B.F.M. Astron. Astrophys. 352, 228 (1999) Transitional YSOs: candidates from flat-spectrum IRAS sources

J/A+A/352/297/  table1.dat  (45Kb)

Stasinska G., Szczerba R. Astron. Astrophys. 352, 297 (1999) The dust content of planetary nebulae: a reappraisal

J/A+A/352/555/  table1.dat  (1.8 Mb)

Allende Prieto C., Lambert D.L. Astron. Astrophys. 352, 555 (1999) Fundamental parameters of nearby stars from the comparison with evolutionary calculations: masses, radii and effective temperatures

J/A+A/352/600/  stars.dat table3.dat table4.dat  (21Kb)

Kaltcheva N.T., Olsen E.H. Astron. Astrophys. 352, 600 (1999) Stroemgren and Hβphotometry of O and B type stars in star-forming regions. I. Canis Major - Puppis - Vela.

J/A+A/352/605/  stars.dat table3.dat table4.dat  (64Kb)

Kaltcheva N.T., Olsen E.H., Clausen J.V. Astron. Astrophys. 352, 605 (1999) Stroemgren and Hβphotometry of O and B type stars in star-forming regions. II. Monoceros OB2, Canis Major OB1 and Collinder 121.

J/A+A/353/77/  refs.dat tablec1.dat  (31Kb)

Teerikorpi P. Astron. Astrophys. 353, 77 (2000) Evidence for the class of the most luminous quasars. II. Variability, polarization, and the gap in the M_{V} distribution.

J/A+A/353/287/  stars.dat table3.dat  (110Kb)

Asplund M., Gustafsson B., Lambert D.L., Rao N.K. Astron. Astrophys. 353, 287 (2000) The R Coronae Borealis stars - atmospheres and abundances

J/A+A/353/528/  table1.dat table1.ps table1.tex  (103Kb)

Ohnaka K., Tsuji T., Aoki W. Astron. Astrophys. 353, 528 (2000) Elemental abundances of carbon, nitrogen, and oxygen in carbon stars

J/A+A/353/958/  catalog.dat  (34Kb)

Scholz R.-D., Irwin M., Ibata R., Jahreiss H., Malkov O.Yu. Astron. Astrophys. 353, 958 (2000) New high-proper motion survey in the Southern sky

J/A+A/353/1044/  table4.dat table4.ps table4.tex  (79Kb)

Hearty T., Neuhaeuser R., Stelzer B., Fernandez M., Alcala J.M., Covino E., Hambaryan V. Astron. Astrophys. 353, 1044 (2000) ROSAT PSPC observations of T Tauri stars in MBM 12

J/A+A/354/169/  table1.dat table2.dat table3.dat table5.dat table6.dat  (37Kb)

Gratton R.G., Sneden C., Carretta E., Bragaglia A. Astron. Astrophys. 354, 169 (2000) Mixing along the red giant branch in metal-poor field stars

J/A+A/354/823/  catalog.dat  (31Kb)

Rozas M., Zurita A., Beckman J.E. Astron. Astrophys. 354, 823 (2000) The ionized gas in the spiral galaxy NGC 3359. I. Photometry.

J/A+A/354/881/  stars.dat table2.dat  (30Kb)

Debernardi Y., Mermilliod J.-C., Carquillat J.-M., Ginestet N. Astron. Astrophys. 354, 881 (2000) Binarity of Am stars in Praesepe and Hyades

J/A+A/354/899/  abund.dat table5.dat table5.tex  (84Kb)

Adelman S.J., Pintado O.I. Astron. Astrophys. 354, 899 (2000) Elemental abundance analyses with complejo astronomico El Leoncito REOSC echelle spectrograms. IV. Extensions of nine previous analyses.

J/A+A/355/L27 See I/259/

J/A+A/355/89/  table3.dat  (38Kb)

Dessauges-Zavadsky M., Pindao M., Maeder A., Kunth D. Astron. Astrophys. 355, 89 (2000) Spectral classification of emission-line galaxies

J/A+A/355/138/  table1.dat  (13Kb)

Bassino L.P., Waldhausen S., Martmnez R.E. Astron. Astrophys. 355, 138 (2000) CCD photometry in the region of NGC 6994: the remains of an old open cluster.

J/A+A/355/171/  table2.dat table4.dat tables.ps tables.tex  (42Kb)

Qian S., Liu Q. Astron. Astrophys. 355, 171 (2000) Orbital period studies of the two contact binaries TZ Bootis and Y Sextantis.

J/A+A/355/398/  table1a.dat table1a.ps table1b.dat table1b.ps tables.tex  (183Kb)

Capitaine N., Guinot B., McCarthy D.D. Astron. Astrophys. 355, 398 (2000) Definition of the Celestial Ephemeris Origin and of UT1 in the International Celestial Reference Frame.

J/A+A/355/651/  table3.dat table4.dat  (27Kb)

Flaccomio E., Micela G., Sciortino S., Damiani F., Favata F., Harnden F.R.Jr, Schachter J. Astron. Astrophys. 355, 651 (2000) HRI observations of PMS stars in NGC 2264.

J/A+A/356/L21/  table1.dat  (9Kb)

Belokon E.T., Babadzhanyants M.K., Pollock J.T. Astron. Astrophys. 356, L21 (2000) ON 231: the testing of 3.8-year period for optical outbursts with multi-peak structure.

J/A+A/356/141/  catalog.dat  (625Kb)

Fabricius C., Makarov V.V. Astron. Astrophys. 356, 141 (2000) Two-colour photometry for 9473 components of close Hipparcos double and multiple stars

J/A+A/356/238/ table2.dat table3.dat table4.dat table5.dat table6.dat table7.dat table8.dat table9.dat (45Kb) Carretta E., Gratton R.G., Sneden C. Astron. Astrophys. 356, 238 (2000) Abundances of light elements in metal-poor stars. III. Data analysis and results

J/A+A/356/445/  table1.dat table2.dat tables.tex  (669Kb)

Brinkmann W., Laurent-Muehleisen S.A., Voges W., Siebert J., Becker R.H., Brotherton M.S., White R.L., Gregg M.D. Astron. Astrophys. 356, 445 (2000) Radio and X-ray bright AGN: the ROSAT - FIRST correlation

J/A+A/356/529/  catalog.dat  (22Kb)

Allen, C., Poveda, A., Herrera, M.A. Astron. Astrophys. 356, 529 (2000) Wide binaries among high-velocity and metal-poor stars

J/A+A/356/541/  table1.dat  (14Kb)

Koehler R., Kunkel M., Leinert C., Zinnecker H. Astron. Astrophys., 356, 541-558 (2000) Multiplicity of X-ray selected T Tauri stars in the Scorpius-Centaurus OB association.

J/A+A/356/590/  table1.dat table2.dat  (11Kb)

Udry S., Mayor M., Naef D., Pepe F., Queloz D., Santos N.C., Burnet M., Confino B., Melo C. Astron. Astrophys. 356, 590 (2000) The CORALIE survey for southern extra-solar planets. II. The short-period planetary companions to HD 75289 and HD 130322.

J/A+A/356/1010/ table2a.dat table2b.dat table2c.dat table2d.dat table2e.dat table2f.dat table4a.dat table4b.dat table4c.dat table4d.dat table4e.dat table4f.dat (306Kb) Wilgenbus D., Cabrit S., Pineau des Forets G., Flower D. Astron. Astrophys. 356, 1010 (2000) The ortho:para-H2 ratio in C- and J- type shock

J/A+A/357/219/  table1.dat table2.dat table3.dat table4.dat tables.ps tables.tex  (67Kb)

Persi P., Marenzi A.M., Olofsson G., Kaas A.A., Nordh L., Huldtgren M., Abergel A., Andre Philippe M., Bontemps S., Boulanger F., Burgdorf M., Casali M.M., Cesarsky C., Copet E., Davies J.I., Falgarone E., Montmerle T., Perault M., Prusti T., Puget J.L., Sibille F. Astron. Astrophys. 357, 219 (2000) ISOCAM observations of the Chamaeleon I dark cloud

J/A+A/357/225/  tablea1.dat tablea2.dat tablea3.dat  (20Kb)

Josselin E., Blommaert J.A.D.L., Groenewegen M.A.T., Omont A., Li F.L. Astron. Astrophys. 357, 225 (2000) Observational investigation of mass loss of M supergiants.

J/A+A/357/367 See I/259/

J/A+A/357/437/  table1.dat table2.dat table3.dat tables.ps tables.tex  (78Kb)

Battinelli P., Capuzzo-Dolcetta R., Hodge P.W., Vicari A., Wyder T.K. Astron. Astrophys. 357, 437 (2000) Large scale star formation in galaxies. I. The spirals NGC 7217, NGC 1058 and UGC 12732. Young star groupings in spirals

J/A+A/357/471/  table2.dat table3.dat table7.dat table9.dat  (310Kb)

Sanner J., Altmann M., Brunzendorf J., Geffert M. Astron. Astrophys. 357, 471 (2000) Photometric and kinematic studies of open star clusters. II. NGC 1960 (M 36) and NGC 2194

J/A+A/357/548/  table2.dat table3.dat table4.dat table5.dat tablea.dat  (36Kb)

Adelman S.J. Astron. Astrophys. 357, 548 (2000) uvby Photometry of the magnetic CP stars HD 36668, 36 Lyncis, HD 86952, and HR 8861

J/A+A/357/572/  tablea.dat  (46Kb)

Cherchneff I., Le Teuff Y.H., Williams P.M., Tielens A.G.G.M. Astron. Astrophys. 357, 572 (2000) Dust formation in carbon-rich Wolf-Rayet stars I. Chemistry of small carbon clusters and silicon species

J/A+A/357/839/  table2.dat table3.dat table4.dat table5.dat tables.ps tables.tex  (90Kb)

Clavel J., Schulz B., Altieri B., Barr P., Claes P., Heras A., Leech K., Metcalfe L., Salama A. Astron. Astrophys. 357, 839 (2000) 2.5-11 micron spectroscopy and imaging of AGNs. Implication for unification schemes

J/A+A/357/871/  field1.dat field2.dat  (210Kb)

Cseresnjes P., Alard C., Guibert J. Astron. Astrophys., 357, 871-880 (2000) Structure of the Sagittarius dwarf galaxy at low Galactic latitudes.

J/A+A/357/1001/  table2.dat table2.ps table2.tex  (58Kb)

Lang W.J., Masheder M.R.W., Dame T.M., Thaddeus P. Astron. Astrophys. 357, 1001 (2000) The λ-Orionis ring in CO.

J/A+A/357/1020/  table1.dat  (74Kb)

Persi P., Tapia M., Roth M. Astron. Astrophys. 357, 1020 (2000) The NGC 6334 IV massive star formation site: A cluster in the making?

J/A+A/358/77/  tablea1.dat tableb1.dat  (51Kb)

Wisotzki L., Christlieb N., Bade N., Beckmann V., Koehler T., Vanelle C., Reimers D. Astron. Astrophys. 358, 77 (2000) The Hamburg/ESO survey for bright QSOs. III. A large flux-limited sample of QSOs.

J/A+A/358/257/ all.ps fig4.ps fig5.ps fig6.ps fig7.ps fig8.ps fig9.ps fig10.ps fig11.ps fig12.ps fig13.ps fig14.ps fig15.ps fig16.ps fig17.ps fig18.ps fig19.ps fig20.ps fig21.ps tab2-3.ps tab2-3.tex table1.ps table2.dat table3.dat (677Kb) Nilsson A., Hjalmarson A., Bergman P., Millar T.J. Astron. Astrophys. 358, 257 (2000) SO and CS observations of molecular clouds. II. Analysis and modelling of the abundance ratios - probing O_{2}/CO with SO/CS?

J/A+A/358/547/  dao.dat ism.dat table1.dat  (300Kb)

Kopacki G. Astron. Astrophys. 358, 547 (2000) Variable stars in the globular cluster M53

J/A+A/358/553/  table5.dat  (25Kb)

Hensberge H., Pavlovski K., Verschueren W. Astron. Astrophys. 358, 553 (2000) The eclipsing binary V578 Mon in the Rosette nebula: age and distance to NGC 2244 using Fourier disentangled component spectra

J/A+A/358/819/  tab5-14.dat  (16Kb)

Weilbacher P.M., Duc P.-A., Fritze-v.Alvensleben U., Martin P., Fricke K.J. Astron. Astrophys. 358, 819 (2000) Tidal dwarf candidates in a sample of interacting galaxies

J/A+A/359/18/  table2.dat  (27Kb)

Molla M., Garcia-Vargas M.L. Astron. Astrophys. 359, 18 (2000) Models for the interpretation of Ca Triplet and the blue spectral indices in elliptical nuclei.

J/A+A/359/113/  notes.dat tablea1.dat tableb1.dat tableb1.ps tableb1.tex  (203Kb)

Grosso N., Montmerle T., Bontemps S., Andre P., Feigelson E.D. Astron. Astrophys. 359, 113 (2000) X-rays and regions of star formation: a combined ROSAT-HRI/near-to-mid IR study of the Rho Oph dark cloud

J/A+A/359/347/  table2.dat table3.dat table4.dat tables.ps tables.tex  (110Kb)

Dutra C.M., Bica E. Astron. Astrophys. 359, 347 (2000) Foreground and background dust in star cluster directions.

J/A+A/359/597/  notes.dat refs.dat table1.ps table1.tex table1a.dat  (137Kb)

Rodriguez E., Lopez-Gonzalez M.J. Astron. Astrophys. 359, 597 (2000) SX Phe stars in globular clusters

J/A+A/359/865/  table4.dat  (22Kb)

Stickel M., Lemke D., Klaas U., Beichman C.A., Rowan-Robinson M., Efstathiou A., Bogun S., Kessler M.F., Richter G. Astron. Astrophys. 359, 865 (2000) The ISOPHOT 170μm Serendipity Survey. I. Compact sources with galaxy associations.

J/A+A/360/439/  tables.dat  (8Kb)

Sarzi M., Corsini E.M., Pizzella A., Vega Beltran J.C., Cappellari M., Funes J.G., Bertola F. Astron. Astrophys. 360, 439 (2000) NGC 4672: A new case of an early-type disk galaxy with an orthogonally decoupled core.

J/A+A/360/472/ table2.dat table3a.dat table3b.dat table5.dat table6.dat table8.dat table9.dat (1.3 Mb) van Leeuwen F., Le Poole R.S., Reijns R.A., Freeman K.C., de Zeeuw P.T. Astron. Astrophys. 360, 472 (2000) A proper motion study of the globular cluster omega Centauri

J/A+A/360/499/  table2.dat table3.dat  (21Kb)

Tautvaisiene G., Edvardsson B., Tuominen I., Ilyin I. Astron. Astrophys. 360, 499 (2000) Chemical composition of evolved stars in the open cluster M 67

J/A+A/360/509/  table3.dat  (16Kb)

Hamdani S., North P., Mowlavi N., Raboud D., Mermilliod J.-C. Astron. Astrophys. 360, 509 (2000) Chemical abundances in seven red giants of NGC 2360 and NGC 2447

J/A+A/360/529/  table3.dat table4.dat table5.dat table6a.dat table6b.dat  (69Kb)

Piatti A.E., Claria J.J., Bica E. Astron. Astrophys. 360, 529 (2000) Photometric and integrated spectral study of the young open clusters Pismis 22, NGC 6178, NGC 6216 and Ruprecht 130

J/A+A/360/1107/  table1.dat  (13Kb)

Deharveng L., Nadeau D., Zavagno A., Caplan J. Astron. Astrophys. 360, 1107 (2000) The stellar content of the compact HII region Sh2-88B

J/A+A/361/245/  table5.dat table6.dat table7.dat table8.dat table9.dat table10.dat  (45Kb)

Aerts C. Astron. Astrophys. 361, 201 (2000) Follow-up photometry of six new βCep stars discovered from the HIPPARCOS mission

J/A+A/361/429/  refs.dat table1.dat table2.dat table3.dat table4.dat table5.dat table6.dat  (29Kb)

Roussel H., Sadat R., Blanchard A. Astron. Astrophys. 361, 429 (2000) The baryon content of groups and clusters of galaxies

J/A+A/361/491/  notes.dat table4.dat table6.dat  (19Kb)

Kahabka P., Puzia T.H., Pietsch W. Astron. Astrophys. 361, 491 (2000) Probing the gas content of the dwarf galaxy NGC 3109 with background X-ray sources.

J/A+A/361/641/  table2.dat table2.ps table2.tex  (199Kb)

Winters J.M., Le Bertre T., Jeong K.S., Helling C., Sedlmayr E. Astron. Astrophys. 361, 641 (2000) A systematic investigation of the mass loss mechanism in dust forming long-period variable stars

J/A+A/361/660/  table1.dat table1.ps table1.tex  (65Kb)

Porras A., Cruz-Gonzalez I., Salas L. Astron. Astrophys. 361, 660 (2000) Young stellar clusters and H_{2} nebulosities in S233IR

J/A+A/361/863/  fig/* table2.dat  (17.0 Mb)

Gavazzi G., Franzetti P., Boselli A., Pierini D., Scodeggio M. Astron. Astrophys. 361, 863 (2000) 1.65 micrometer (H-band) surface photometry of galaxies. V. Profile decomposition of 1157 galaxies.

J/A+A/361/929/  table1.dat table1.ps table1.tex table2.dat table2.ps table2.tex table3.dat  (2.4 Mb)

Meibom S. Astron. Astrophys. 361, 929 (2000) New Color-magnitude Diagram for the Open Cluster IC 4651. Deep Stroemgren uvby CCD photometry in a large field.

J/A+A/361/945/  table1.dat  (24Kb)

Lapasset E., Claria J.J., Mermilliod J.-C. Astron. Astrophys. 361, 945 (2000) UBV photometric study and basic parameters of the southern open cluster NGC 2539

J/A+A/361/991/  table2.dat  (5Kb)

Wade G.A., Debernardi Y., Mathys G., Bohlender D.A., Hill G.M., Landstreet J.D. Astron. Astrophys. 361, 991 (2000) An analysis of the Ap binary HD 81009.

J/A+A/361/1005/  table1.dat  (19Kb)

Richter G.A., Greiner J. Astron. Astrophys. 361, 1005 (2000) S10947 Aql = RX J2009.8+1557: A probable RS CVn star which sometimes stops its eclipses

J/A+A/361/1011/  position.dat table1.dat table2.dat table3.dat  (31Kb)

Pasquini L., De Medeiros J.R., Girardi L. Astron. Astrophys. 361, 1011 (2000) Ca II activity and rotation in F-K evolved stars.

J/A+A/361/1023/ hbtracks.dat isojcmag.dat isostmag.dat isovega.dat mstracks.dat sspjcmag.dat sspstmag.dat sspvega.dat (46.6 Mb) Salasnich B., Girardi L., Weiss A., Chiosi C. Astron. Astrophys., 361, 1023 (2000) Evolutionary tracks and isochrones for α-enhanced stars.

J/A+A/361/1079/ table1.dat table2.dat table5a.dat table5b.dat table6a.dat table6b.dat table7.dat table8.dat table9a.dat table9b.dat table10.dat table11.dat table12.dat table13.dat table14.dat (43Kb) Zinchenko I., Henkel C., Mao R.Q. Astron. Astrophys. 361, 1079 (2000) HNCO in massive galactic dense cores

J/A+A/361/1143/  table4.dat table5.dat table6.dat  (37Kb)

Teixeira R., Ducourant C., Sartori M.J., Camargo J.I.B., Perie J.P., Lepine J.R.D., Benevides-Soares P. Astron. Astrophys. 361, 1143 (2000) Proper motions of pre-main sequence stars in southern star-forming regions.

J/A+A/362/199/  table4.dat  (26Kb)

Felli M., Comoretto G., Testi L., Omont A., , Schuller F. Astron. Astrophys. 362, 199 (2000) The search for YSOs from ISOGAL data. Application to the l=+45 field

J/A+A/362/215/  table2a.dat table2b.dat  (104Kb)

Schultheis M., Ganesh S., Glass I.S., Omont A., Ortiz R., Simon G., van Loon J., Alard C., Blommaert J.A.D.L., Borsenberger J., Fouque P., Habing H.J.

Astron. Astrophys. 362, 215 (2000) DENIS and ISOGAL properties of variable star candidates in the Galactic Bulge

J/A+A/362/223/  table1.dat  (47Kb)

Jetsu L., Hackman T., Hall D.S., Henry G.W., Kokko M., You J. Astron. Astrophys. 362, 223 (2000) Time series analysis of V815 Herculis photometry between 1984 and 1998

J/A+A/362/519/  refs.dat tablea1.dat  (26Kb)

De Breuck C., Rottgering H., Miley G., van Breugel W., Best P. Astron. Astrophys. 362, 519 (2000) A statistical study of emission lines from high redshift radio galaxies

J/A+A/362/673/  b-v.dat b.dat u-b.dat u.dat v.dat  (43Kb)

Kallrath J., Strassmeier K.G. Astron. Astrophys. 362, 673 (2000) The BF Aurigae system. A close binary at the onset of mass transfer.

J/A+A/362/799/  table1.dat  (22Kb)

Giommi P., Perri M., Fiore F. Astron. Astrophys. 362, 799 (2000) The BeppoSAX 2-10keV survey.

J/A+A/363/L1/  lmcrrd.dat  (21Kb)

Kovacs G. Astron. Astrophys. 363, L1 (2000) The distance modulus of the Large Magellanic Cloud based on double-mode RR Lyrae stars.

J/A+A/363/29/  table8.dat table9.dat  (14Kb)

Antonello E., Fugazza D., Mantegazza L., Bossi M., Covino S. Astron. Astrophys. 363, 29 (2000) Variable stars in nearby galaxies. III. White light observations of Field B of IC 1613.

J/A+A/363/62/  tablea1.dat  (61Kb)

Rousseau J., Paturel G., Vauglin I., Schroeder A., de Batz B., Borsenberger J., Epchtein N., Fouque P., Kimeswenger S., Lacombe F., Le Bertre T., Mamon G., Rouan D., Simon G., Tiphene D. Astron. Astrophys., 363, 62-68 (2000) Some noteworthy DENIS galaxies near the galactic plane.

J/A+A/363/141/  table1.dat table2.dat  (207Kb)

Reich W., Fuerst E., Reich P., Kothes R., Brinkmann W., Siebert J. Astron. Astrophys. 363, 141 (2000) Radio-loud active galaxies in the northern ROSAT All-Sky Survey. IV. Properties of sources stronger than 100 mJy at 5 GHz.

J/A+A/363/239/  table2.dat  (17Kb)

De Medeiros J.R., do Nascimento J.D.Jr, Sankarankutty S., Costa J.M., Maia M.R.G. Astron. Astrophys. 363, 239 (2000) Rotation and lithium in single giant stars.

J/A+A/363/440/  table2.dat  (9Kb)

Liang H., Lemonon L., Valtchanov I., Pierre M., Soucail G. Astron. Astrophys. 363, 440 (2000) Probing the gravitational potential of a nearby lensing cluster Abell 2104.

J/A+A/363/517/  table1.dat table1.ps table1.tex  (56Kb)

Mobasher B., Mazzei P. Astron. Astrophys. 363, 517 (2000) The effect of dust on photometric redshift measurement: a self-consistent technique.

J/A+A/363/887/  table1.dat  (16Kb)

Dallacasa D., Stanghellini C., Centonza M., Fanti R. Astron. Astrophys. 363, 887 (2000) High frequency peakers. I. The bright sample.

J/A+A/363/991 See I/260/

J/A+A/363/1040/  fig/* spectra.dat table2.dat table3a.dat  (829Kb)

Friedrich S., Koester D., Christlieb N., Reimers D., Wisotzki L. Astron. Astrophys. 363, 1040 (2000) Cool helium-rich white dwarfs from the Hamburg/ESO survey.

J/A+A/363/1051/ colors.dat farb.f farby he0/* he1/* hgamma/* mod.f models.dat models/* profiles.dat table1.dat table1.tex (7.3 Mb) Zboril M. Astron. Astrophys. 363, 1051 (2000) Properties of He-rich stars. III. Model atmospheres and diagnostic tools.

J/A+A/363/1081/ atlas.dat atlasb.dat atlasbco.dat atlasco.dat original.tar phoenibc.dat phoenix.dat phoenixb.dat phoenixc.dat sun.dat sunmon.dat vega.dat vegamon.dat (32.7 Mb) Claret A. Astron. Astrophys. 363, 1081 (2000) A new non-linear limb-darkening law for LTE stellar atmosphere models: Calculations for -5.0 log[M/H] +1, 2000K T_{eff} 50000K at several surface gravities

J/A+A/364/102/  stars.dat table4.dat table5.dat  (46Kb)

Kinman T., Castelli F., Cacciari C., Bragaglia A., Harmer D., Valdes F. Astron. Astrophys. 364, 102 (2000) A spectroscopic study of field BHB star candidates

J/A+A/364/205/  stars.dat table2.dat table3.dat tables.tex ubvmag.dat ymag.dat  (57Kb)

Cutispoto G., Pastori L., Guerrero A., Tagliaferri G., Messina S., Rodono M., de Medeiros J.R. Astron. Astrophys. 364, 205 (2000) Photometric and spectroscopic studies of cool stars discovered in EXOSAT X-ray images. IV. The northern hemisphere sample.

J/A+A/364/237/  abund.dat tablea1.dat tablea1.ps tablea1.tex  (125Kb)

Albayrak B. Astron. Astrophys. 364, 237 (2000) A spectral analysis of Deneb (A2 Iae).

J/A+A/364/391/  table1.dat table4.dat table5.dat  (19Kb)

Lobanov, A.P., Krichbaum, T.P., Graham, D.A., Witzel, A., Kraus, A., Zensus, J.A., Britzen, S., Greve, A., Grewing, M. Astron. Astrophys. 364, 391 (2000) 86GHz VLBI survey of compact radio sources

J/A+A/364/455/  tab982.dat tab1433.dat tab11284.dat tab12578.dat  (4.9 Mb)

Larsen S.S., Clausen J.V., Storm J. Astron. Astrophys. 364, 455 (2000) Reddenings and metallicities in the LMC and SMC from Stroemgren CCD photometry.

J/A+A/364/517/  table1.dat table1.ps table1.tex  (43Kb)

Rodighiero G., Granato G.L., Franceschini A., Fasano G., Silva L. Astron. Astrophys. 364, 517 (2000) Spiral and irregular galaxies in the Hubble Deep Field North. Comparison with early types and implications for the global SFR density.

J/A+A/364/665/  refs.dat stars.dat tab4-10.dat tab11-14.dat table15.dat  (32Kb)

Segransan D., Delfosse X., Forveille T., Beuzit J.-L., Udry S., Perrier C., Mayor M. Astron. Astrophys. 364, 665 (2000) Accurate masses of very low mass stars: III. 16 new or improved masses

J/A+A/364/674/  stars.dat table7.dat  (198Kb)

Castilho B.V., Gregorio-Hetem J., Spite F., Barbuy B., Spite M. Astron. Astrophys. 364, 674 (2000) Detailed analysis of a sample of Li-rich giants

J/A+A/364/689/  table1.dat  (8Kb)

Manfroid J., Mathys G. Astron. Astrophys. 364, 689 (2000) Variations of the Ap star HDE 318107

J/A+A/364/712/  table5.dat table6.dat  (32Kb)

Penprase B.E., Rhodes J.D., Harris E.L. Astron. Astrophys. 364, 712 (2000) Optical observations of the Draco molecular cloud. I. Catalog of B and V magnitudes for selected areas.

J/A+A/365/1/  table2.dat table4.dat tables.ps tables.tex  (31Kb)

Matthews L.D., van Driel W., Monnier-Ragaigne D. Astron. Astrophys. 365, 1 (2001) HI observations of giant low surface brightness galaxies

J/A+A/365/90/  stars.dat table4.dat table5.dat table6.dat table7.dat  (21Kb)

Folha D.F.M., Emerson J.P. Astron. Astrophys. 365, 90 (2001) Near infrared hydrogen lines as diagnostic of accretion and winds in T Tauri Stars

J/A+A/365/110/  stars.dat table1.dat table2.dat  (29Kb)

Eiroa C., Garzon F., Alberdi A., de Winter D., Ferlet R., Grady C.A., Cameron A., Davies J.K., Deeg H.J., Harris A.W., Horne K., Merin B., Miranda L.F., Montesinos B., Mora A., Oudmaijer R., Palacios J., Penny A., Quirrenbach A., Rauer H., Schneider J., Solano E., Tsapras Y., Wesselius P.R. Astron. Astrophys. 365, 110 (2001) EXPORT: Near-IR observations of Vega-type and pre-main sequence stars

J/A+A/365/222 See I/266/

J/A+A/365/266/ table1.dat table2.dat table3.dat table4.dat table5.dat table6.dat table7.dat table8.dat table9.dat table10.dat table11.dat table12.dat table13.dat table14.dat table15.dat table16.dat table17.dat (64Kb) Tully J.A., Chidichimo M.C. Astron. Astrophys. 365, 266 (2001) Oscillator strengths between fine structure levels of Fe XXIII.

J/A+A/365/268/  table1.dat table2.dat table3.dat table4.dat table5.dat tables.ps tables.tex  (98Kb)

Bautista M. Astron. Astrophys. 365, 268 (2001) Atomic data from the Iron Project. XLVII. Electron impact excitation of Ni III

J/A+A/365/275 See VIII/66/

J/A+A/365/370/  table2.dat table3.dat  (34Kb)

Perez-Gonzalez P. G., Gallego J., Zamorano J., Gil de Paz A. Astron. Astrophys. 365, 370 (2001) Optical photometry of the UCM Lists I and II. II. B band surface photometry and morphological discussion.

J/A+A/365/409/  table3.dat table4.dat  (9Kb)

Oyabu S., Kawara K., Tsuzuki Y., Sofue Y., Sato Y., Okuda H., Taniguchi Y., Shibai H., Gabriel C., Hasegawa T. and Nishihara E. Astron. Astrophys. 365, 409 (2001) ISO continuum observations of quasars at z=1-4. I. Spectral energy distributions of quasars from UV to far-infrared

J/A+A/365/545/  table2.dat table3.dat table4.dat tables.ps tables.tex  (115Kb)

Habing H.J., Dominik C., Jourdain de Muizon M., Laureijs R.J., Kessler M.F., Leech K., Metcalfe L., Siebenmorgen R., Trams N., Bouchet P. Astron. Astrophys. 365, 545 (2001) Incidence and survival of remnant disks around main-sequence stars

J/A+A/365/656/  table1.dat table2.dat tables.ps tables.tex  (72Kb)

Popovic L.C., Milovanovic N., Dimitrijevic M.S. Astron. Astrophys. 365, 656 (2001) The electron-impact broadening effect in hot star atmospheres: The case of singly- and doubly-ionized zirconium.

J/A+A/366/68/  fits/* table1.dat table3.dat  (2.2 Mb)

Prugniel P., Maubon G., Simien F. Astron. Astrophys. 366, 68 (2001) The formation of galaxy bulges: Spectrophotometric constraints

J/A+A/366/100/  table2.dat  (32Kb)

Simon V., Shugarov S., Marsakova V.I. Astron. Astrophys. 366, 100 (2001) Long-term color variations of the peculiar X-ray binary V Sagittae.

J/A+A/366/428/  table1.dat table2.dat tables.ps tables.tex  (104Kb)

Karachentsev I.D., Karachentseva V.E., Huchtmeier W.K. Astron. Astrophys. 366, 428 (2001) HI observations of nearby galaxies. III. More dwarf galaxies in the northern sky.

J/A+A/366/481/  table2.dat table2.ps table2.tex  (263Kb)

Sevenster M.N., van Langevelde H.J., Moody R.A., Chapman J.M., Habing H.J., Killeen N.E.B. Astron. Astrophys. 366, 481 (2001) The ATCA/VLA OH 1612 MHz survey. III. Observations of the Northern Galactic Plane.

J/A+A/366/498/  table2.dat  (17Kb)

Chandar R., Bianchi L., Ford H. C.

Astron. Astrophys. 366, 498 (2001) Star clusters in M33. IV. A new survey from deep HST images.

J/A+A/366/508/ notes1.dat notes2.dat notes3.dat refs.dat stars.dat table1.dat table2.dat table3.dat table16.dat (79Kb) van Genderen A.M. Astron. Astrophys. 366, 508 (2001) S Doradus variables in the Galaxy and the Magellanic Clouds

J/A+A/366/538 See VI/102/

J/A+A/366/565/  appenxa.dat  (17Kb)

Kokkotas K.D., Ruoff J. Astron. Astrophys. 366, 565 (2001) Radial oscillations of relativistic stars

J/A+A/366/771/  table3.dat table4.dat table5.dat table6.dat table7.dat table8.dat  (37Kb)

Kniazev A.Y., Engels D., Pustilnik S.A., Ugryumov A.V., Kniazeva T.F., Pramsky A.G., Brosch N., Hagen H.-J., Hopp U., Izotov Y.I., Lipovetsky V.A., Masegosa J., Marquez I., Martin J.-M. Astron. Astrophys. 366, 771 (2001) The Hamburg/SAO Survey for Emission-Line Galaxies. IV. The Fourth List of 119 Galaxies

J/A+A/366/868/  table2.dat  (13Kb)

Docobo J.A., Tamazian V.S., Balega Y.Y., Blanco J., Maximov A.F., Vasyuk V.A. Astron. Astrophys. 366, 868 (2001) Binary star speckle measurements at Calar Alto. I.

J/A+A/366/1003 See III/238/

J/A+A/367/86/  table9.dat table10.dat  (17Kb)

Smartt S.J., Venn K.A., Dufton P.L., Lennon D.J., Rolleston W.R.J., Keenan F.P. Astron. Astrophys. 367, 86 (2001) Chemical abundances in the inner 5 kpc of the Galactic disk

J/A+A/367/111/  notes.dat tablea1.dat tablea2.dat  (54Kb)

de Bruijne J.H.J., Hoogerwerf R., de Zeeuw P.T. Astron. Astrophys. 367, 111 (2001) A Hipparcos study of the Hyades open cluster: improved colour-absolute magnitude and Hertzsprung-Russell diagrams.

J/A+A/367/199/  table2.dat table3.dat  (20Kb)

Skopal A., Teodorani M., Errico L., Vittone A.A., Ikeda Y., Tamura S. Astron. Astrophys. 367, 199 (2001) A photometric and spectroscopic study of the eclipsing symbiotic binary AX Persei

J/A+A/367/236/  table3.dat  (33Kb)

Lehmann H., Harmanec P., Aerts C, Bozic H., Eenens P., Hildebrandt G., Holmgren D., Mathias P., Scholz G., Slechta M., Yang S. Astron. Astrophys. 367, 236 (2001) A new analysis of the radial-velocity variations of the eclipsing and spectroscopic binary EN Lacertae

J/A+A/367/297/  table2.dat  (50Kb)

Adelman S.J. Astron. Astrophys. 367, 297 (2001) Hipparcos photometry: The least variable stars.

J/A+A/367/498/  table1.dat table2.dat  (15Kb)

Magrini L., Cardwell A., Corradi R., Mampaso A., Perinotto M. Astron. Astrophys. 367, 498 (2001) Accurate Positions of Candidate Planetary Nebulae in M33

J/A+A/367/521 See II/224/

J/A+A/367/597/  table6.ps table6.tex table6a.dat table6b.dat  (202Kb)

Adelman S.J., Gulliver A.F., Rayle K.E. Astron. Astrophys. 367, 597 (2001) Elemental abundance analyses with DAO spectrograms. XXIV. The Mercury Manganese stars υHer, ϕHer, and HR 7018

J/A+A/367/725/  table2.dat table3.dat  (28Kb)

Wroblewski H., Costa E. Astron. Astrophys. 367, 725 (2001) High proper motion stars with declinations between -30^{∘}and -40^{∘}, and right ascensions between 00h and 10h

40m.

J/A+A/367/759/  table3.dat table4.dat  (14Kb)

Mantegazza L., Antonello E., Fugazza D., Bossi M., Covino S. Astron. Astrophys. 367, 759 (2001) Variable stars in nearby galaxies. IV. Fields C and D of IC 1613.

J/A+A/367/859/  table4.ps table4.tex table4a.dat table4b.dat  (188Kb)

Lopez-Garcia Z., Adelman S.J., Pintado O.I. Astron. Astrophys. 367, 859 (2001) Elemental abundance studies of CP stars. III: The magnetic CP stars αScl and HD 170973

J/A+A/367/865/  table2.dat  (8Kb)

Prieur J.L., Oblak E., Lampens P., Kurpinska-Winiarska M., Aristidi E., Koechlin L., Ruymakers G. Astron. Astrophys. 367, 865 (2001) Speckle observations of binary systems measured by Hipparcos

J/A+A/367/910/ stars.dat table1.dat table3.dat table4.dat table5.dat tables.tex ubv.dat ubvri.dat (84Kb) Cutispoto G., Messina S., Rodono M. Astron. Astrophys. 367, 910 (2001) Long-term monitoring of active stars. IX. Photometry collected in 1993.

J/A+A/367/1049/  table3.dat  (109Kb)

Zhang H.W., Zhao G., Hu J.Y. Astron. Astrophys. 367, 1049 (2001) A catalogue of emission lines in spectra of Comet C/1995 O1(Hale-Bopp)

J/A+A/368/16/  table5.dat table5.ps table5.tex  (45Kb)

Moellenhoff C., Heidt J. Astron. Astrophys. 368, 16 (2001) Surface photometry of spiral galaxies in NIR: structural parameters of disks and bulges

J/A+A/368/225/  table2.dat table3.dat table4.dat table5.dat tablea.dat  (35Kb)

Adelman S.J. Astron. Astrophys. 368, 225 (2001) uvby Photometry of the CP stars HR 5341, HD 142070, HR 6967, and HR 8434.

J/A+A/368/250/  table11.dat table12.dat tables.ps tables.tex  (45Kb)

Arellano Ferro A., Giridhar S., Mathias P. Astron. Astrophys. 368, 250 (2001) Atmospheric abundances in post-AGB candidates of intermediate temperature

J/A+A/368/267/  table4.dat  (30Kb)

Maccarone M.C., Mineo T., Preite-Martinez A. Astron. Astrophys. 368, 267 (2001) Spatially resolved spectroscopy of Cassiopeia A with MECS on board Beppo-SAX.

J/A+A/368/335 See II/272/

J/A+A/368/797/  figs1/* figs2/* seyfert1.dat seyfert2.dat  (12.0 Mb)

Pfefferkorn F., Boller T., Rafanelli P. Astron. Astrophys. 368, 797 (2001) Soft X-ray properties of a spectroscopically selected sample of interacting and isolated Seyfert galaxies.

J/A+A/368/835/  table2.dat table3.dat  (18Kb)

Sidoli L., Belloni T., Mereghetti S. Astron. Astrophys. 368, 835 (2001) A catalogue of soft X-ray sources in the Galactic Center region.

J/A+A/368/845/  refs.dat table2.dat  (163Kb)

Valdettaro R., Palla F., Brand J., Cesaroni R., Comoretto G., Di Franco S., Felli M., Natale E., Palagi F., Panella D., Tofani G. Astron. Astrophys. 368, 845 (2001) The Arcetri Catalog of H_{2}O maser sources: Update 2000

J/A+A/368/912/  appen.ps appen.tex appen1.dat appen2.dat  (329Kb)

Yudin R.V. Astron. Astrophys. 368, 912 (2001) Statistical analysis of intrinsic polarization, IR excess and projected rotational velocity distributions of classical Be stars.

J/A+A/368/932/  plot table.dat  (189Kb)

Zola S., Ogloza W.

Astron. Astrophys. 368, 932 (2001) Accretion disk in the binary system V367 Cygni

J/A+A/368/1006/  refs.dat table1.dat  (150Kb)

Chen P.S., Szczerba R., Kwok S., Volk K. Astron. Astrophys. 368, 1006 (2001) Properties of OH/IR stars with IRAS LRS spectra

J/A+A/368/1021 See V/106/

J/A+A/369/57/  table4.dat table5.dat tables.ps tables.tex  (62Kb)

Santos-Lleo M., Clavel J., Schulz B., Altieri B., Barr P., Alloin D. Berlind P., Bertram R., Crenshaw D, Edelson R.A., Giveon U., Horne K., Huchra J.P., Kaspi S., Kriss G.A., Krolik J.H., Malkan M.A., Malkov Yu.F., Netzer H., O’Brien P.T., Peterson B.M., Pogge R.W., Pronik V.I., Qian B.-C., Reichert G.A., Rodriguez-Pascual P.M., Sergeev S.G., Tao J., Tokarz S., Wagner R.M., Wamsteker W., Wilkes B.J. Astron. Astrophys. 369, 57 (2001) Monitoring of the optical and 2.5-11.7 micrometer spectrum and mid-IR imaging of the Seyfert 1 galaxy Mkn 279 with ISO.

J/A+A/369/178/  stars.dat table10.dat  (56Kb)

Bergeat J., Knapik A., Rutily B. Astron. Astrophys. 369, 178 (2001) The effective temperatures of carbon-rich stars

J/A+A/369/249/  stars.dat tablea1.dat  (26Kb)

Woitas J., Koehler R., Leinert C. Astron. Astrophys. 369, 249 (2001) Orbital motion in T Tauri binary systems

J/A+A/369/380/  table3.dat table3.ps table3b.dat table3b.ps  (232Kb)

Fanti C., Pozzi F., Dallacasa D., Fanti R., Gregorini L., Stanghellini C., Vigotti M. Astron. Astrophys. 369, 380 (2001) Multi-frequency VLA observations of a new sample of CSS/GPS radio sources

J/A+A/369/511/  table5.dat table7.dat table8.dat table12.dat table13.dat table14.dat  (1.2 Mb)

Sanner J., Brunzendorf J., Will J.-M., Geffert M. Astron. Astrophys. 369, 511 (2001) Photometric and Kinematic Studies of open star clusters. III. NGC 4103, NGC 5281, and NGC 4755 (kappa Cru)

J/A+A/369/527/  table5.dat table5.ps table5.tex  (149Kb)

van Houten C.J. Astron. Astrophys. 369, 527 (2001) Observations of OB-stars at the former Leiden Southern Station

J/A+A/369/552/  table1.dat  (9Kb)

Salaman M., Morlet G., Gili R. Astron. Astrophys. 369, 552 (2001) Nice Observatory CCD measurements of visual double stars (3rd series)

J/A+A/369/758/  table1.dat  (23Kb)

Fan J.H., Qian B.C., Tao J. Astron. Astrophys. 369, 758 (2001) Optical observations of BL Lacertae from 1997 to 1999

J/A+A/369/812/  table3.dat table4.dat  (20Kb)

Rejkuba M. Astron. Astrophys. 369, 812 (2001) Deep VLT search for globular clusters in NGC 5128 : Color-magnitude diagrams and globular cluster luminosity function.

J/A+A/369/862/  dao-i.dat dao-v.dat ism-i.dat ism-v.dat table1.dat  (648Kb)

Kopacki G. Astron. Astrophys. 369, 862 (2001) Variable stars in the globular cluster M 92.

J/A+A/369/932/  table3.dat table5.dat  (439Kb)

Kontizas E., Dapergolas A., Morgan D.H., Kontizas M. Astron. Astrophys. 369, 932 (2001) A Catalogue of carbon stars in the LMC.

J/A+A/369/960/  table2.dat  (11Kb)

Manimanis V.N., Niarchos P.G. Astron. Astrophys. 369, 960 (2001) A Photometric study of the near-contact system RU Ursae Minoris

J/A+A/369/1048 See III/218/

J/A+A/370/78/  table3.dat table4.dat  (24Kb)

Mingaliev M.G., Stolyarov V.A., Davies R.D., Melhuish S.J., Bursov N.A., Zhekanis G.V. Astron. Astrophys. 370, 78 (2001) Observations of the bright radio sources in the North Celestial Pole region at the RATAN-600 radio telescope.

J/A+A/370/409/  table1.dat table2.dat  (15Kb)

Lara L., Cotton W.D., Feretti L., Giovannini G., Marcaide J.M., Marquez I., Venturi T. Astron. Astrophys. 370, 409 (2001) A new sample of large angular size radio galaxies. I. The radio data.

J/A+A/370/426/  table5.dat  (7Kb)

Dessauges-Zavadsky M., D’Odorico S., McMahon R.G., Molaro P., Ledoux C., Peroux C., Storrie-Lombardi L.J. Astron. Astrophys. 370, 426 (2001) UVES observations of a damped Lyαsystem at z_{abs} = 4.466 towards the quasar APM BR J0307-4945.

J/A+A/370/468/ survey.dat table1.dat table1.ps table1.tex table5.txt table5a.dat table5b.dat table6.dat table7.dat table8.dat (474Kb) Torres D.F., Romero G.E., Combi J.A., Benaglia P., Andernach H., Punsly B. Astron. Astrophys. 370, 468 (2001) A variability analysis of low-latitude unidentified gamma-ray sources

J/A+A/370/496/  table1.dat  (9Kb)

Kiss L.L., Skiff B.A. Astron. Astrophys. 370, 496 (2001) The radial pulsation of AI Aurigae

J/A+A/370/503/  table1.dat  (28Kb)

Munari U., Jurdana-Sepic R., Moro D. Astron. Astrophys. 370, 503 (2001) Symbiotic stars on Asiago archive plates

J/A+A/370/765/  table1.dat table2.dat table3.dat table4.dat table5.dat  (59Kb)

Verheijen M.A.W., Sancisi R. Astron. Astrophys. 370, 765 (2001) The Ursa Major cluster of galaxies. IV. HI synthesis observations.

J/A+A/370/900/  table1.dat  (13Kb)

Vignali C., Comastri A., Fiore F., La Franca F. Astron. Astrophys. 370, 900 (2001) The BeppoSAX High Energy Large Area Survey. IV. On the soft X-ray properties of the hard X-ray-selected HELLAS sources.

J/A+A/370/931/  table2.dat table3.dat table4.dat table5.dat  (92Kb)

Piatti A.E., Claria J.J. Astron. Astrophys. 370, 931 (2001) On the stellar content of the open clusters Melotte 105, Hogg 15, Pismis 21 and Ruprecht 140

J/A+A/370/974/  table2.dat  (13Kb)

Duemmler R., Aarum V. Astron. Astrophys. 370, 974 (2001) The radial velocities of the RS CVn star UX Ari. A triple system with a binary on the same line of sight.

J/A+A/370/1071/  table2.dat table3.dat tables.ps tables.tex  (111Kb)

Binello A.M., Landi E., Mason H.E., Storey P.J., Brosius J.W. Astron. Astrophys. 370, 1071 (2001) A comparison between theoretical and solar Fe XII UV line intensity ratios.

J/A+A/371/L5/  table1.dat table2.dat  (151Kb)

Maitzen H.M., Paunzen E., Pintado O.I. Astron. Astrophys. 371, L5 (2001) On the detection of the first extragalactic classical chemically peculiar stars.

J/A+A/371/79/  table1a.dat table2a.dat table3a.dat  (35Kb)

Dietrich M., Bender C.F., Bergmann D.J., Bills T.E., Bochkarev N.G., Burenkov A.N., Gaskell C.M., Gutzmer D.D., Grove R., Hiller M.E., Huchra J.P., Klimek E.S., Lund C., Merkulova N.I., Pebley S., Poulsen M.A., Pronik V.I., Sergeev S.G., Sergeev E.A., Shapovalova A.I., Vlasyuk V.V., Wilkes B. Astron. Astrophys. 371, 79 (2001) A spectroscopic and photometric study of short-timescale variability in NGC 5548

J/A+A/371/497/  table1.dat table3.dat  (13Kb)

Pietrzynski G., Gieren W., Fouque P., Pont F. Astron. Astrophys. 371, 497 (2001) A catalog of OB associations in the spiral galaxy NGC 300.

J/A+A/371/579/  clusbv.dat clusv.dat clusvi.dat fieldv.dat stars.dat table1.dat  (251Kb)

Kovacs G., Walker A.R. Astron. Astrophys. 371, 579 (2001) Empirical relations for cluster RR Lyrae stars revisited

J/A+A/371/833/  table3.dat table4.dat  (27Kb)

Lehmann I., Hasinger G., Schmidt M., Giacconi R., Truemper J., Zamorani G., Gunn J.E., Pozzetti L., Schneider D.P., Stanke T., Szokoly G., Thompson D., Wilson G. Astron. Astrophys. 371, 833 (2001) The ROSAT Deep Survey: VI. X-ray sources and Optical identifications of the Ultra Deep Survey.

J/A+A/371/908/  stars.dat table1.dat  (63Kb)

Vazquez R.A., Baume G. Astron. Astrophys. 371, 908 (2001) The open cluster Havlen-Mofat No. 1 revisited

J/A+A/371/943/  parcat.dat  (17Kb)

Chen Y.Q., Nissen P.E., Benoni T.,G.Zhao. Astron. Astrophys. 371, 943 (2001) Lithium abundances for 185 main-sequence stars - Galactic evolution and stellar depletion of lithium

J/A+A/371/986/  table3.dat table4.dat table5.dat  (53Kb)

Poretti E. Astron. Astrophys. 371, 986 (2001) Fourier decomposition and frequency analysis of the pulsating stars with P 1d in the OGLE database. I. Monoperiodic Delta Scuti, RRc and RRab variables. Separation criteria and particularities.

J/A+A/371/1078/  table5.ps table5.tex table5a.dat table5b.dat  (135Kb)

Adelman S.J., Caliskan H., Kocer D., Kablan H., Yuce K., Engin S. Astron. Astrophys. 371, 1078 (2001) Elemental abundance analyses with DAO Spectrograms. XXV. The superficially normal B and A stars alpha Draconis, tau Herculus, gamma Lyrae, and HR 7926.

J/A+A/372/L41/  table1.dat  (11Kb)

Motte F., Andre P., Ward-Thompson D., Bontemps S. Astron. Astrophys. 372, L41 (2001) A SCUBA survey of the NGC 2068/2071 protoclusters.

J/A+A/372/71/  table2.dat table3.dat table4.dat  (86Kb)

Durgapal A.K., Pandey A.K., Mohan V. Astron. Astrophys. 372, 71 (2001) Multicolor CCD photometry of old open clusters: King 5 and Be 20

J/A+A/372/85/  stars.dat table3.dat  (32Kb)

Ramdani A., Jorissen A. Astron. Astrophys. 372, 85 (2001) Evolution of the dust mass loss with luminosity along the giant branch of the globular cluster 47 Tuc.

J/A+A/372/95/  table2.dat  (15Kb)

Kaltcheva N.T., Gredel R., Fabricius C. Astron. Astrophys. 372, 95 (2001) The association surrounding NGC 2439

J/A+A/372/145/  table2.dat table2a.ps table2b.ps table2c.ps table2d.ps table3.dat table3.ps  (140Kb)

Henden A., Munari U Astron. Astrophys. 372, 145 (2001) UBV(RI)c comparison sequences for symbiotic stars. II.

J/A+A/372/173/  table1.dat table2.dat table3.dat table4.dat table5.dat  (42Kb)

Bontemps S., Andre P., Kaas A.A., Nordh L., Olofsson G., Huldtgren M., Abergel A., Blommaert J., Boulanger F., Burgdorf M., Cesarsky C.J., Cesarsky D., Copet E., Davies J., Falgarone E., Lagache G., Montmerle T., Perault M., Persi P., Prusti T., Puget J.L., Sibille F.

Astron. Astrophys. 372, 173 (2001) ISOCAM observations of the rho Ophiuchi cloud: Luminosity and mass functions of the pre-main sequence embedded cluster.

J/A+A/372/245/  table1.dat  (31Kb)

Chen P.S. Astron. Astrophys. 372, 245 (2001) Infrared properties of barium stars

J/A+A/372/276/  table1.dat  (33Kb)

Berdyugin A.V., Teerikorpi P., Haikala L., Hanski M.O., Knude J., Markkanen T. Astron. Astrophys. 372, 276 (2001) Interstellar polarization at high galactic latitudes from distant stars. IV. A catalog of polarization data for the North Galactic Pole Area

J/A+A/372/364/ table7.dat table8.dat table9.dat table10.dat table11.dat table12.dat tables.ps tables.tex (71Kb) Dole H., Gispert R., Lagache G., Puget J.-L., Bouchet F.R., Cesarsky C., Ciliegi P., Clements D.L., Dennefeld M., Desert F.-X., Elbaz B., Franceschini A., Guiderdon B., Harwit M., Lemke D., Moorwood A.F.M., Oliver S., Reach W.T., Rowan-Robinson M., Stickel M. Astron. Astrophys. 372, 364 (2001) FIRBACK: III. Catalog, source counts, and cosmological implications of the 170 microns ISO Deep Survey

J/A+A/372/391/  table4.dat  (100Kb)

Czoske O., Kneib J.-P., Soucail G., Bridges T.J., Mellier Y., Cuillandre J.-C. Astron. Astrophys. 372, 391 (2001) A wide-field spectroscopic survey of the cluster of galaxies C10024+1654. I. The catalogue.

J/A+A/372/477/  table2a.dat table2b.dat table8.dat table9.dat table14a.dat table14b.dat  (54Kb)

Marco A., Bernabeu G. Astron. Astrophys. 372, 477 (2001) Photometric study of the double cluster h + χPersei

J/A+A/372/588/  table1.dat  (62Kb)

Rodriguez E., Garcia J.M., Martin S., Claret A. Astron. Astrophys. 372, 588 (2001) uvby light curves of the eclipsing binary system V2154 Cyg

J/A+A/372/862/  stars.dat table2.dat table3.dat table4.dat tables.ps tables.tex  (51Kb)

Randich S., Pallavicini R., Meola G., Stauffer J.R., Balachandran S.C. Astron. Astrophys. 372, 862 (2001) Membership, lithium, and metallicity in the young open clusters IC 2602 and IC 2391: enlarging the sample

J/A+A/372/879/  table2.dat  (65Kb)

Carraro G., Hassan, S.M., Ortolani S., Vallenari A. Astron. Astrophys. 372, 879 (2001) An optical and near infrared study of the old open cluster NGC 2141

J/A+A/372/1078/  osc.dat readups.f table3.dat ups.dat  (1.2 Mb)

Butler K., Zeippen C.J. Astron. Astrophys. 372, 1078 (2001) Atomic data from the IRON Project. XLIX. Electron impact excitation of Fe XX.

J/A+A/372/1083/  osc.dat readups.f table2.dat ups.dat  (1.4 Mb)

Butler K., Zeippen C.J. Astron. Astrophys. 372, 1083 (2001) Atomic data from the IRON Project. L: Electron impact excitation of Fe XIX.

J/A+A/373/24/  notes.dat table2.dat table3.dat tablea1.dat  (37Kb)

Pustilnik S.A., Kniazev A.Y., Lipovetsky V.A., Ugryumov A.V. Astron. Astrophys. 373, 24 (2001) Environment status of blue compact galaxies and trigger of star formation.

J/A+A/373/63/  table5.dat table6.dat table7.dat table8.dat  (133Kb)

Supper R., Hasinger G., Lewin W.H.G., Magnier E.A., van Paradijs J., Pietsch W., Read A.M., Truemper J. Astron. Astrophys. 373, 63 (2001) The second ROSAT PSPC survey of M31 and the complete ROSAT PSPC source list.

J/A+A/373/153/  table4.dat table5.dat table6.dat table7.dat table8.dat  (30Kb)

Paunzen E., Maitzen H.M. Astron. Astrophys. 373, 153 (2001)

CCD photometric search for peculiar stars in open clusters. II. NGC 2489, NGC 2567, NGC 2658, NGC 5281 and NGC 6208

J/A+A/373/159 See III/221A/

J/A+A/373/438/  table2.dat  (37Kb)

Haberl F., Pietsch W. Astron. Astrophys. 373, 438 (2001) The X-ray view of M 33 after ROSAT.

J/A+A/373/536/  table1.dat table2.dat tables.ps tables.tex  (68Kb)

van de Steene G. C., Jacoby G. Astron. Astrophys. 373, 536 (2001) Radio observations of new galactic bulge planetary nebulae.

J/A+A/373/576/  table3.dat  (141Kb)

Piquard S., Halbwachs J.-L., Fabricius C., Geckeler R., Soubiran C., Wicenec A. Astron. Astrophys. 373, 576 (2001) Variable stars in the Tycho photometric observations: I. Detection.

J/A+A/373/608/ table1.dat table2.dat table3.dat table4.dat table5.dat table6.dat table7.dat table8.dat (42Kb) Dufton P.L., Smartt S.J., Hambly N.C. Astron. Astrophys. 373, 608 (2001) A UKST survey of blue objects towards the Galactic Centre - seven additional fields.

J/A+A/373/625/  table2.dat  (65Kb)

Paunzen E., Duffee B., Heiter U., Kuschnig R., Weiss W.W. Astron. Astrophys. 373, 625 (2001) A spectroscopic survey for lambda Bootis stars. II. The observational data.

J/A+A/373/730/ tablea1.dat tablea2.dat tablea3.dat tablea4.dat tablea5.dat tablea6.dat tableb1.dat (29Kb) Gonzalez-Riestra R., Cassatella A., Wamsteker W. Astron. Astrophys. 373, 730 (2001) The INES System IV: The IUE Absolute Flux Scale

J/A+A/373/757/  tablea1.dat tablea2.dat tablea3.dat  (40Kb)

Kim T.-S., Cristia S., D’Odorico S. Astron. Astrophys. 373, 757 (2001) The Lyman-Alpha Forest at 1.5 z 4

J/A+A/373/899/ fig/* table1.dat table2.dat table3.dat table4.dat table5.dat table6.dat table7.dat table8.dat table9.dat table10.dat table11.dat table12.dat table13.dat table14.dat table15.dat table16.dat table17.dat table18.dat table19.dat table20.dat table21.dat table22.dat table23.dat table24.dat table25.dat table26.dat table27.dat table28.dat table29.dat table30.dat table31.dat table32.dat table33.dat table34.dat table35.dat table36.dat table37.dat table38.dat table39.dat (805Kb) Kaluzny J., Thompson I.B. Astron. Astrophys. 373, 899 (2001) Clusters AgeS Experiment. New variable stars in the globular cluster M22

J/A+A/373/950/  table4.dat  (54Kb)

Galis R., Hric L., Niarchos P. Astron. Astrophys. 373, 950 (2001) KW Persei - a near-contact system?

J/A+A/373/987/  legend.dat xsec/*  (1.3 Mb)

Zboril M., Berrington K.A. Astron. Astrophys. 373, 987 (2001) Non-LTE gallium abundance in HgMn stars.

J/A+A/373/1032/  refs.dat table1.dat table1.ps table1.tex  (123Kb)

Siodmiak N., Tylenda R. Astron. Astrophys. 373, 1032 (2001) An analysis of the observed radio emission from planetary nebulae.

J/A+A/374/92 See VII/224/

J/A+A/374/132/  stars.dat table3.dat table4.dat table5.dat  (77Kb)

Lampens P., Oblak E., Duval D., Chareton M. Astron. Astrophys. 374, 132 (2001) CCD photometry and astrometry for visual double and multiple stars of the HIPPARCOS catalogue. III.

CCD photometry and differential astrometry for 253 southern ”intermediate” systems

J/A+A/374/164/  data.dat  (18Kb)

Zola S., Niarchos P., Manimanis V., Dapergolas A. Astron. Astrophys. 374, 164 (2001) A Photometric study of BH Cas

J/A+A/374/204/  table1.dat  (24Kb)

Debernardi Y., North P. Astron. Astrophys. 374, 204 (2001) Eclipsing binaries with candidate CP stars. II. Parameters of the system V392 Carinae.

J/A+A/374/227/  table1.dat table4.dat  (24Kb)

Tokovinin A.A., Gorynya N.G. Astron. Astrophys. 374, 227 (2001) New spectroscopic components in multiple systems. IV.

J/A+A/374/265/  table4a.dat table4b.dat table4c.dat  (1.9 Mb)

Wade G.A., Bagnulo S., Kochukhov O., Landstreet J.D., Piskunov N., Stift M.J. Astron. Astrophys. 374, 265 (2001) LTE spectrum synthesis in magnetic stellar atmospheres. The interagreement of three independent polarised radiative transfer codes.

J/A+A/374/394/  galaxies.dat table5.dat table6.dat  (63Kb)

Vega Beltran J.C., Pizzella A., Corsini E.M., Funes J.G., Zeilinger W.W., Beckman J., Bertola F. Astron. Astrophys. 374, 394 (2001) Kinematic properties of gas and stars in 20 disc galaxies

J/A+A/374/504/  table2.dat  (2.0 Mb)

Pandey, A.K., Nilakshi, Ogura, K., Sagar, R., Tarusawa, K. Astron. Astrophys. 374, 504 (2001) NGC 7654: An interesting cluster to study star formation history

J/A+A/374/638/  observ.dat  (31Kb)

Djurasevic G., Zakirov M., Eshankulova M., Erkapic S. Astron. Astrophys. 374, 638 (2001) A photometric study of the massive binary RY Sct

J/A+A/374/712/  j2 0706.dat j2 0709.dat k1 0704.dat n4 0704.dat n4 0706.dat  (9Kb)

Szabo Gy.M., Csak B., Sarneczky K., Kiss L.L. Astron. Astrophys. 374, 712 (2001) Photometric observations of distant active comets

J/A+A/374/861/  tablea1.dat tablea2.dat tablea3.dat tablea4.dat  (29Kb)

Schoenmakers A.P., de Bruyn A.G., Rottgering H.J.A., van der Laan H. Astron. Astrophys. 374, 861 (2001) A new sample of giant radio galaxies from the WENSS survey

J/A+A/374/907/  table3.dat table4.dat table5.dat table6.dat table7.dat tables.ps tables.tex  (176Kb)

Ugryumov A.V., Engels D., Kniazev A.Y., Green R.F., Izotov Y.I., Hopp U., Pustilnik S.A., Pramsky A.G., Kniazeva T.F., Brosch N., Hagen H.-J., Lipovetsky V.A., Masegosa J., Marquez I., Martin J.-M Astron. Astrophys. 374, 907 (2001) The Hamburg/SAO Survey for Emission-Line Galaxies. V. The Fifth List of 161 Galaxies.

J/A+A/374/968/  annex.dat annex.ps annex.tex  (273Kb)

Mennessier M.O., Mowlavi N., Alvarez R., Luri X. Astron. Astrophys. 374, 968 (2001) Long Period Variable Stars: galactic populations and infrared luminosity calibrations.

J/A+A/374/1049/  table2a.dat table2b.dat table2c.dat table2d.dat table2e.dat  (113Kb)

Korhonen H., Berdyugina S.V., Tuominen I., Andersen M.I., Piironen J., Strassmeier K.G., Grankin K.N., Kaasalainen S., Karttunen H., Mel’nikov S.Yu., Shevchenko V.S., Trisoglio M., Virtanen J. Astron. Astrophys. 374, 1049 (2001) Study of FK Comae Berenices III. Photometry for the years 1993–2001

J/A+A/374/1056/  table4a.dat table4b.dat  (10Kb)

Arentoft T., Sterken C., Handler G., Freyhammer L.M., Bruch A., Niarchos P., Gazeas K., Manimanis V., Van Cauteren P., Poretti E., Dawson D.W., Liu Z.L., Zhou A.Y., Du B.T., Shobbrook R.R., Garrido R., Fried R., Akan M .C., Ibanoglu C., Evren S., Tas G., Johnson D., Blake C., Kurtz D .W. Astron. Astrophys. 374, 1056 (2001) V1162 Ori: A multiperiodic delta Scuti star with variable period and amplitude.

J/A+A/374/1085/  tablea.dat tablea.ps tablea.tex  (122Kb)

Cassatella A., Altamore A., Badiali M., Cardini D. Astron. Astrophys. 374, 1085 (2001) On the Wilson-Bappu relationship in the Mg II K line

J/A+A/375/L27/  table1.dat table2.dat table3.dat  (10Kb)

Naef D., Latham D.W., Mayor M., Mazeh T., Beuzit J.L., Drukier G.A., Perrier-Bellet C., Queloz D., Sivan J.P., Torres G., Udry S., Zucker S. Astron. Astrophys. 375, L27 (2001) HD 80606b, a planet on an extremely elongated orbit

J/A+A/375/1/  cdfkcat.dat hdfkcat.dat  (112Kb)

Saracco P., Giallongo S., Cristiani S., D’Odorico S., Fontana A., Iovino A., Poli F., Vanzella E. Astron. Astrophys. 375, 1 (2001) Deep near-IR observations of the Chandra Deep Field and of the HDF South. Color and number counts.

J/A+A/375/30/  table1.dat  (18Kb)

Mermilliod J.-C., Claria J.J., Andersen J., Piatti A.E., Mayor M. Astron. Astrophys. 375, 30 (2001) Red giants in open clusters: IX. NGC 2324, 2818, 3590 and 6259

J/A+A/375/130/  stars.dat table2.dat table3.dat  (31Kb)

Covino E., Melo C., Alcala J.M., Torres G., Fernandez M., Frasca A., Paldino R. Astron. Astrophys. 375, 130 (2001) New low-mass pre-main sequence spectroscopic binaries in Orion

J/A+A/375/205/  table1.dat table2.dat table3.dat table4.dat table5.dat table6.dat  (18Kb)

Naef D., Mayor M., Pepe F., Queloz D., Santos N.C., Udry S., Burnet M. Astron. Astrophys. 375, 205 (2001) The CORALIE survey for southern extrasolar planets. V. 3 new extrasolar planets

J/A+A/375/275/  table3.dat table4.dat table5.dat  (233Kb)

Baudrand A., Bec-Borsenberger A., Borsenberger J., Barucci M.A. Astron. Astrophys. 375, 275 (2001) Asteroidal I, J, K magnitudes recovered in the DENIS Survey: first release

J/A+A/375/285/ 699time.dat 1865ph.dat 1866time.dat cv3time.dat gk137ph.dat jd6ph.dat nd43time.dat nmph.dat rh27time.dat (61Kb) Szabo G.M., Csak B., Sarneczky K., Kiss L.L. Astron. Astrophys. 375, 285 (2001) Photometric observations of 9 Near-Earth Objects.

J/A+A/375/308/  regions.dat vmcc.dat  (3.2 Mb)

Camargo J.I.B., Teixeira R., Benevides-Soares P., Ducourant C. Astron. Astrophys., 375, 308-318 (2001) Extension of the ICRF for selected areas down to the 15th V magnitude

J/A+A/375/366/  hesfhlcs.dat  (54Kb)

Christlieb NN., Green P.J., Wisotzki L., Reimers D. Astron. Astrophys. 375, 366 (2001) The stellar content of the Hamburg/ESO survey. II. A large, homogeneously-selected sample of high latitude carbon stars.

J/A+A/375/527/ stars.dat tableb2.dat tableb3.dat tableb4.dat tableb5.dat tables.ps tables.tex (71Kb) Kerschbaum F., Lebzelter T., Lazaro C. Astron. Astrophys. 375, 527 (2001) Multi-colour light variation of AGB stars observed with ISO

J/A+A/375/539/  table2.dat  (332Kb)

Comeron F., Torra J. Astron. Astrophys. 375, 539 (2001) Near-infrared imaging of compact HII regions in Cygnus X

J/A+A/375/591/  tablea1.dat tablea1.pdf tablea1.tex  (446Kb)

Curdt W., Brekke P., Feldman U., Wilhelm K., Dwivedi B.N., Schuhle U., Lemaire P. Astron. Astrophys. 375, 591 (2001) The SUMER Spectral Atlas of Solar Disk Features.

J/A+A/375/614/  table3.dat table4.dat  (156Kb)

Noel F. Astron. Astrophys. 375, 614 (2001)

Astrometric observations of the Sun at Santiago: 1998-2000.

J/A+A/375/661/  table2.dat table2.ps table2.tex  (119Kb)

Gontier A.-M., Le Bail K., Feissel M., Eubanks T.M. Astron. Astrophys. 375, 661 (2001) Stability of the extragalactic VLBI reference frame

J/A+A/375/670/  table2.dat table2.ps table2.tex  (66Kb)

Rizvanov N., Dautov I., Shaimukhametov R. Astron. Astrophys. 375, 670 (2001) The comparative accuracy of photographic observations of radio stars observed at the Engelhardt astronomical observatory

J/A+A/375/739/  refs.dat table5.dat  (44Kb)

Donato D., Ghisellini G., Tagliaferri G., Fossati G. Astron. Astrophys. 375, 739 (2001) Hard X-ray properties of blazars.

J/A+A/375/863/  table3.dat  (1.1 Mb)

Jeffries R.D., Thurston M.R., Hambly N.C. Astron. Astrophys. 375, 863 (2001) Photometry and membership for low mass stars in the young open cluster NGC 2516

J/A+A/375/909/  lcurves/* table1.dat table2.dat table3.dat table4.dat  (150Kb)

Cseresnjes P. Astron. Astrophys. 375, 909 (2001) RR Lyrae stars in the Sagittarius dwarf galaxy: Period analysis

J/A+A/375/989/  table1.dat  (10Kb)

Bouvier J., Duchene G., Mermilliod J.-C., Simon T. Astron. Astrophys. 375, 989 (2001) The formation and evolution of binary systems. III. Low-mass binaries in the Praesepe cluster.

J/A+A/376/28/  tablea1.dat  (12Kb)

Rollinde E., Petitjean P., Pichon C. Astron. Astrophys. 376, 28 (2001) Physical properties and small-scale structure of the Lyman-alpha forest: Inversion of the HE 1122-1648 UVES spectrum.

J/A+A/376/144/  table2.dat table3.dat  (179Kb)

Pigulski A., Kopacki G., Kolaczkowski Z. Astron. Astrophys. 376, 144 (2001) The young open cluster NGC 663 and its Be stars.

J/A+A/376/158/  table1.dat table1.ps table1.tex  (45Kb)

Albayrak B., Demircan O., Djurasevic G., Erkapic S., Ak H. Astron. Astrophys. 376, 158 (2001) A light and period study of SV Camelopardalis

J/A+A/376/325 See I/272/

J/A+A/376/333/  table1a.dat table1b.dat  (22Kb)

Sharma A.K., Chandra S. Astron. Astrophys. 376, 333 (2001) Suggestions for an interstellar cyclopropene search.

J/A+A/376/356/  table1.dat table3.dat  (275Kb)

Chandra S., Sharma A.K. Astron. Astrophys. 376, 356 (2001) Collisional rates for vib-rotational transitions in diatomic molecules.

J/A+A/376/441/  clusters.dat ori.tar pm/* table1.dat table2.dat  (1.5 Mb)

Dias W.S., Lepine J.R.D., Alessi B.S. Astron. Astrophys. 376, 441 (2001) Proper motions of open clusters within 1 kpc based on the TYCHO2 Catalogue.

J/A+A/376/561/  table2.dat v1.dat v2.dat v3.dat v4.dat v5.dat v6.dat v7.dat  (55Kb)

Kiss L.L., Szabo G.M., Sziladi K., Furesz G., Sarneczky K., Csak B. Astron. Astrophys. 376, 561 (2001) A variable star survey of the open cluster M 37.

J/A+A/376/745/  table4.dat  (13Kb)

Petrov G., Seggewiss W., Dieball A., Kovachev B.

Astron. Astrophys. 376, 745 (2001) CCD standards for U and I in the open cluster NGC 7790

J/A+A/376/775/  table2.dat table4.dat table5.dat tables.ps tables.tex  (64Kb)

Shapovalova A.I., Burenkov A.N., Carrasco L., Chavushyan V.H., Doroshenko V.T., Dumont A.M., Lyuty V.M., Valdes J.R., Vlasuyk V.V., Bochkarev N.G., Collin S., Legrand F., Mikhailov V.P., Spiridonova O.I., Kurtanidze O., Nikolashvili M.G. Astron. Astrophys. 376, 775 (2001) Intermediate resolution H-beta spectroscopy and photometric monitoring of 3C 390.3. I. Further evidence of a nuclear accretion disk.

J/A+A/376/861/  figs/* figsj.tar lb.tar lbj.tar maps.dat  (27.2 Mb)

Reich P., Testori J.C., Reich W. Astron. Astrophys. 376, 861 (2001) A radio continuum survey of the southern sky at 1420 MHz.

J/A+A/376/907/  catalog.dat  (16Kb)

Persi P., Marenzi A.R., Gomez M., Olofsson G. Astron. Astrophys. 376, 907 (2001) Sub-arcsec resolution near-infrared images of the Cederblad 110 region.

J/A+A/376/982/ figb1p1.ps figb1p2.ps figb1p3.ps figc1p1.ps figc1p2.ps figc1p3.ps figc1p4.ps figc1p5.ps figc1p6.ps figc1p8.ps stars.dat tablea1.dat tablea2.dat (1.4 Mb) Woitas J., Leinert C., Koehler R. Astron. Astrophys. 376, 982 (2001) Mass ratios of the components in T Tauri binary systems and implications for multiple star formation

J/A+A/376/997/  table4.dat  (11Kb)

Le Bertre T., Matsuura M., Winters J.M., Murakami H., Yamamura I., Freund M., Tanaka M. Astron. Astrophys. 376, 997 (2001) Galactic mass-losing AGB stars probed with the IRTS. I.

J/A+A/376/1106/  table4.dat table4.ps table4.tex  (75Kb)

Charro E., Martin I. Astron. Astrophys. 376, 1106 (2001) Relativistic quantum defect orbital study of forbidden transitions in Co^{16+}

J/A+A/376/1113/  gline.dat rline.dat taba1-3.dat taba4-6.dat  (933Kb)

Porquet D., Mewe R., Dubau J., Raassen A.J.J., Kaastra J.S. Astron. Astrophys. 376, 1113 (2001) Line ratios for helium-like ions: Applications to collision-dominated plasmas.

J/A+A/376/1123/  table1.dat  (13Kb)

Adraou A., Chini R., Albrecht M., Lemke R., Shaver P.A., Nyman L.-A., Booth R.S. Astron. Astrophys. 376, 1123 (2001) Pointing sources for southern submillimeter telescopes

J/A+A/377/123/ hd008071.dat hd013357.dat hd030101.dat hd033334.dat hd200466.dat hd219542.dat stars.dat (57Kb) Gratton R.G., Bonanno G., Claudi R.U., Cosentino R., Desidera S., Lucatello S., Scuderi S. Astron. Astrophys. 377, 123 (2001) Non-interacting main-sequence binaries with different chemical compositions: evidences of infall of rocky material?

J/A+A/377/361/  table2.dat table2.ps table2.tex table3.dat table3.ps table3.tex  (139Kb)

Pickering J.C., Johansson S., Smith P.L. Astron. Astrophys. 377, 361 (2001) The FERRUM Project: Branching ratios and atomic transition probabilities of Fe II transitions from the 3d^{6}(a^{3}F)4p subconfiguration in the visible to VUV spectral region

J/A+A/377/428/  table1.dat  (9Kb)

Sodre L.Jr., Proust D., Capelato H.V., Lima Neto G.B., Cuevas H., Quintana H., Fouque P. Astron. Astrophys. 377, 428 (2001) The cluster of galaxies Abell 970

J/A+A/377/442/  table2.dat  (11Kb)

Wolf C., Meisenheimer K., Dye S., Kleinheinrich M., Rix H.-W., Wisotzki L. Astron. Astrophys. 377, 442 (2001) Deep BVR photometry of the Chandra Deep Field South from the COMBO-17 survey

J/A+A/377/801/  table1.dat table2.dat tables.ps tables.tex  (141Kb)

Huchtmeier W.K., Karachentsev I.D., Karachentseva V.E Astron. Astrophys. 377, 801 (2001) HI observations of nearby galaxies. IV. More dwarf galaxies in the southern sky.

J/A+A/377/911/  table1.dat  (284Kb)

Feltzing S., Holmberg J., Hurley J.R. Astron. Astrophys. 377, 911 (2001) The solar neighbourhood age-metallicity relation. Does it exist?

J/A+A/377/925/  tablea1.dat tablea2.dat tablea3.dat tables.ps tables.tex  (127Kb)

Barziv O., Kaper L., van Kerkwijk M.H., Telting J.H., van Paradijs J. Astron. Astrophys. 377, 925 (2001) The mass of the neutron star in Vela X-1.

J/A+A/377/945/  tablea2.dat tablea3.dat  (26Kb)

Cioni M.-R.L., Marquette J.-B., Loup C., Azzopardi M., Habing H.J., Lasserre T., Lesquoy E. Astron. Astrophys. 377, 945 (2001) Variability and spectral classification of LMC giants: results from DENIS and EROS

J/A+A/378/30/  images/* table1.dat  (82.6 Mb)

Zimmermann H.-U., Boller T., Doebereiner S., Pietsch W. Astron. Astrophys. 378, 30 (2001) Galaxy identifications in the ROSAT All-Sky Survey.

J/A+A/378/113/  catalog.tex notes.dat periods.dat stars.dat  (224Kb)

Renson P., Catalano F.A. Astron. Astrophys. 378, 113 (2001) Fifth supplement to the catalogue of observed periods of Ap stars.

J/A+A/378/116/  refs.dat table1.dat table2.dat table6.dat  (27Kb)

Mora A., Merin B., Solano E., Montesinos B., de Winter D., Eiroa C., Ferlet R., Grady C.A., Davies J.K., Miranda L.F., Oudmaijer R.D., Palacios J., Quirrenbach A., Harris A.W., Rauer H., Cameron A., Deeg H.J., Garzon F., Penny A., Schneider J., Tsapras Y., Wesselius P.R.. Astron. Astrophys. 378, 116 (2001) EXPORT: spectral classification and projected rotational velocities of Vega-type and pre-main sequence stars.

J/A+A/378/556/  tablea1.dat tablea1.ps  (136Kb)

Koester D., Napiwotzki R., Christlieb N., Drechsel H., Hagen H.J., Heber U., Homeier D., Karl C., Leibundgut B., Moehler S., Nelemans G., Reimers D., Reinecke E-M., Renzini A., Yungelson L. Astron. Astrophys. 378, 556 (2001) High-resolution UVES/VLT spectra of white dwarfs for the ESO SN Ia progenitor survey (SPY). I.

J/A+A/378/861/ fit/* stars.dat table2.dat table3.dat table4.dat table5.dat table6.dat table7.dat tables.tex (4.2 Mb) Chauville J., Zorec J., Ballereau D., Morrell N., Cidale L., Garcia A. Astron. Astrophys. 378, 861 (2001) High and intermediate-resolution spectroscopy of Be stars. An atlas of Hγ, He I 4471 and Mg II 4481 lines.

J/A+A/378/954/  table2.dat  (38Kb)

Gili R., Bonneau D. Astron. Astrophys. 378, 954 (2001) CCD Measurements of visual double stars made with the 74cm and 50cm refractors of the Nice Observatory (2nd series)

J/A+A/379/21/  table1.dat table2.dat  (1.8 Mb)

Zanichelli A., Vigotti M., Scaramella R., Grueff G., Vettolani G. Astron. Astrophys. 379, 21 (2001) Radio-optically selected clusters of galaxies. I. The radiogalaxy sample.

J/A+A/379/35/  table1.dat table2.dat  (14Kb)

Zanichelli A., Scaramella R., Vettolani G., Vigotti M., Bardelli S., Zamorani G. Astron. Astrophys. 379, 35 (2001) Radio-optically selected clusters of galaxies. II. The cluster sample.

J/A+A/379/54/  table2.dat table3.dat table4.dat tablea1.dat tables.ps tables.tex  (103Kb)

Hernandez-Toledo H.M., Puerari I. Astron. Astrophys. 379, 54 (2001) BVRI surface photometry of (S+S) binary galaxies. I. The data.

J/A+A/379/90/  table1.dat  (12Kb)

Magrini L., Corradi R.L.M., Perinotto M., Mampaso A.

Astron. Astrophys. 379, 90 (2001) New candidate planetary nebulae in M 81

J/A+A/379/136/  coll228.dat ngc3114.dat tables.dat vdbh99.dat  (420Kb)

Carraro G., Patat F. Astron. Astrophys. 379, 136 (2001) Star clusters in the Carina complex: UBVRI photometry of NGC 3114, Collinder 228 and vdB-Hagen 99.

J/A+A/379/162/  table2.dat table5.dat table7.dat table9.dat  (34Kb)

Gerbaldi M., Faraggiana R., Balin N. Astron. Astrophys. 379, 162 (2001) Binary systems with post-T Tauri secondaries

J/A+A/379/305/  table3.dat  (25Kb)

Alvarez R., Jorissen A., Plez B., Gillet D., Fokin A., Dedecker M. Astron. Astrophys. 379, 305 (2001) Envelope tomography of long-period variable stars. III. Line doubling frequency among Mira stars.

J/A+A/379/453/  table2.dat table3.dat table4.dat table5.dat table6.dat  (130Kb)

Piatti A.E., Claria J.J. Astron. Astrophys. 379, 453 (2001) On the nature of five open cluster candidates projected close to the Galactic centre

J/A+A/379/564/  stars.dat tablea1.dat  (73Kb)

Oudmaijer R.D., Palacios J., Eiroa C., Davies J.K., de Winter D., Ferlet R., Garzon F., Grady C.A., Cameron A., Deeg H.J., Harris A.W., Horne K., Merin B., Miranda L.F., Montesinos B., Mora A., Penny A., Quirrenbach A., Rauer H., Schneider J., Solano E., Tsapras Y., Wesselius. P. Astron. Astrophys. 379, 564 (2001) EXPORT: Optical photometry and polarimetry of Vega-type and pre-main sequence stars

J/A+A/379/740/ b.dat b filt.dat i.dat i filt.dat r.dat r filt.dat u.dat u filt.dat up.dat up filt.dat v.dat v filt.dat (40.4 Mb) Arnouts S., Vandame B., Benoist C., Groenewegen M.A.T., da Costa L., Schirmer M., Mignani R.P., Slijkhuis R. Astron. Astrophys. 379, 740 (2001) ESO imaging survey. Deep public survey: Multi-color optical data for the Chandra Deep Field South.

J/A+A/379/755/  table3.dat table3.ps table3.tex  (207Kb)

Venturi T., Dallacasa D., Orfei A., Bondi M., Fanti C., Gregorini L., Mantovani F., Stanghellini C., Trigilio C., Umana G. Astron. Astrophys. 379, 755 (2001) Radio monitoring of a sample of X- and gamma-ray loud blazers.

J/A+A/379/798/  table2.dat  (11Kb)

Pascual S., Gallego J., Aragon-Salamanca A., Zamorano J. Astron. Astrophys. 379, 798 (2001) Hαemitting galaxies and the star formation rate density at z = 0.24.

J/A+A/379/976/  stars.dat table1.dat table4.dat table5.dat  (26Kb)

Montes D., Lopez-Santiago J., Fernandez-Figueroa M.J., Galvez M.C. Astron. Astrophys. 379, 976 (2001) Chromospheric activity, lithium and radial velocities of single late-type stars possible members of young moving groups.

J/A+A/379/999/  table1.dat table2.dat  (9Kb)

Santos N.C., Mayor M., Naef D., Pepe F., Queloz D., Udry S., Burnet M. Astron. Astrophys. 379, 999 (2001) The CORALIE survey for southern extrasolar planets. VI: New long period giant planets around HD 28185 and HD 213240

J/A+A/380/238/  table1.dat table2.dat  (31Kb)

Shatsky N. Astron. Astrophys. 380, 238 (2001) The long-period companions of multiple stars tend to have moderate eccentricities

J/A+A/380/258/  table.dat  (10Kb)

Ostrov P.G. Astron. Astrophys. 380, 258 (2001) Fundamental parameters for the eclipsing binary AzV 73 in the Small Magellanic Cloud.

J/A+A/380/300/  fluxes.dat nebulae.dat tables.ps tables.tex  (70Kb)

Escudero A.V., Costa R.D.D Astron. Astrophys. 380, 300 (2001) Abundances of recently discovered planetary nebulae towards the Galactic bulge

J/A+A/380/347/  table2.dat table3.dat  (15Kb)

Delsanti A.C., Boehnhardt H., Barrera L., Meech K.J, Sekiguchi T., Hainaut O.R. Astron. Astrophys. 380, 347 (2001) BVRI photometry of 27 Kuiper Belt Objects with ESO/ Very Large Telescope

J/A+A/380/441/  table1.dat table3.dat  (1.4 Mb)

Woudt P.A., Kraan-Korteweg R.C. Astron. Astrophys. 380, 441 (2001) A catalogue of galaxies behind the Southern Milky Way. II. The Crux and Great Attractor regions (l=289-338^{∘})

J/A+A/380/478/  ground.dat hst.dat  (1.2 Mb)

Marconi G., Andreuzzi G., Pulone L., Cassisi S., Testa V., Buonanno R. Astron. Astrophys. 380, 478 (2001) Mass segregation of different populations inside the cluster NGC 6101.

J/A+A/380/578/  stars.dat table1.dat  (21Kb)

Tautvaisiene G., Edvardsson B., Tuominen I., Ilyin I. Astron. Astrophys. 380, 578 (2001) Chemical composition of red horizontal branch stars in the thick disk of the Galaxy

J/A+A/380/590/  table45.dat  (9Kb)

Phan-Bao N., Guibert J., Crifo F., Delfosse X., Forveille T., Borsenberger J., Epchtein N., Fouque P., Simon G. Astron. Astrophys., 380, 590-598 (2001) New neighbours: IV. 30 DENIS late-M dwarfs between 15 and 30 parsecs.

J/A+A/380/609/  table1.dat table2.dat table3.dat table4.dat  (164Kb)

de Winter D., van den Ancker M.E., Maira A., The P.S., Tjin A Djie H.R.E., Redondo I., Eiroa C., Molster F.J. Astron. Astrophys. 380, 609 (2001) A photometric catalogue of southern emission-line stars.

J/A+A/380/665/  table1.dat table1.tex  (48Kb)

Wu Y., Wu J., Wang J. Astron. Astrophys. 380, 665 (2001) A search for massive dense cores with ^{13}CO J=1-0 lines.

J/A+A/380/727/  vienne.dat  (1.4 Mb)

Vienne A., Thuillot W., Veiga C.H., Arlot J.-E., Vieira Martins R. Astron. Astrophys. 380, 727 (2001) Saturnian satellite observations made in Brazil during the 1995 opposition with an astrometric analysis.

J/A+A/381/32/ table1.dat table2.dat table12.ps table12.tex tablea1.dat tablea1.tex tablea1 1.ps tablea1 2.ps (501Kb) Andrievsky S.M., Kovtyukh V.V., Luck R.E., Lepine J.R.D., Bersier D., Maciel W.J., Barbuy B., Klochkova V.G., Panchuk V.E., Karpischek R.U. Astron. Astrophys. 381, 32 (2002) Using Cepheids to determine the Galactic abundance gradient. I. The solar neighbourhood.

J/A+A/381/65/  table4.dat  (975Kb)

Nilakshi N., Sagar R. Astron. Astrophys. 381, 65 (2002) A comprehensive study of the rich open star cluster NGC 2099 based on deep BVI CCD observations

J/A+A/381/105/  table4.dat  (25Kb)

Royer F., Gerbaldi M., Faraggiana R., Gomez A.E. Astron. Astrophys., 381, 105 (2002) Rotational velocities of A-type stars. I. Measurement of vsini in the southern hemisphere.

J/A+A/381/420/  tables.dat  (11Kb)

Adami C., Mazure A. Astron. Astrophys. 381, 420 (2002) Group analysis in the SSRS2 catalog.

J/A+A/381/446/ table1.dat table2a.dat table2b.dat table2c.dat table2d.dat table2e.dat table2f.dat table2g.dat table2h.dat table2i.dat table2j.dat (94Kb)

Madsen S., Dravins D., Lindegren L. Astron. Astrophys. 381, 446 (2002) Astrometric Radial Velocities. III. Hipparcos Measurements of nearby star clusters and associations.

J/A+A/381/464/  table5.dat  (47Kb)

Wu Z.Y., Tian K.P., Balaguer-Nunez L., Jordi C., Zhao J.L., Guibert J. Astron. Astrophys. 381, 464 (2002) Determination of proper motions and membership of the open star cluster NGC 2548

J/A+A/381/500/  table2.dat table5.dat table6.dat  (14Kb)

Koch A., Edvardsson B. Astron. Astrophys. 381, 500 (2002) Eu abundances in F and G disk stars.

J/A+A/381/507/  table1.dat table2.dat  (18Kb)

Kotnik-Karuza D., Friedjung M., Selvelli P.L. Astron. Astrophys. 381, 507 (2002) The region of Fe II emission line formation of the symbiotic nova RR Tel.

J/A+A/381/757/  refs.dat rtable1.dat table1.dat table1.ps table1.tex  (657Kb)

Liu F.K., Zhang Y.H. Astron. Astrophys. 381, 757 (2002) A new list of extra-galactic radio jets

J/A+A/381/884/  table3.dat  (125Kb)

Giorgi E.E., Vazquez R.A., Baume G., Seggewiss W., Will J.-M. Astron. Astrophys. 381, 884 (2002) NGC 2571: An intermediate-age open cluster with a white dwarf candidate.

J/A+A/381/941/  table2.dat  (11Kb)

Heydari-Malayeri M., Charmandaris V., Deharveng L., Meynadier F., Rosa M.R., Schaerer D., Zinnecker H. Astron. Astrophys. 381, 941 (2002) Resolving the compact HII regions in N160A with HST.

J/A+A/382/60/  table1.dat table2.dat table3.dat table4.dat  (16Kb)

Duc P.-A., Poggianti B.M., Fadda D., Elbaz D., Flores H., Chanial P., Franceschini A., Moorwood A., Cesarsky C. Astron. Astrophys. 382, 60 (2002) Hidden star-formation in the cluster of galaxies Abell 1689.

J/A+A/382/92/  table1.dat table2.dat table3.dat  (26Kb)

Shatsky N., Tokovinin A. Astron. Astrophys. 382, 92 (2002) The mass ratio distribution of B-type visual binaries in the Sco OB2 association

J/A+A/382/118/  notes.dat table1.dat table2.dat table6.dat  (176Kb)

Tokovinin A.A., Smekhov M.G. Astron. Astrophys. 382, 118 (2002) Statistics of spectroscopic sub-systems in visual multiple stars

J/A+A/382/184/ stars.dat table4.dat table5.dat table6.dat table7.dat table8.dat table9.dat table10.dat table11.dat table12.dat table13.dat table14.dat table15.dat table16.dat table17.dat table18.dat table19.dat table20.dat (59Kb) Molster F.J., Waters L.B.F.M., Tielens A.G.G.M., Barlow M.J. Astron. Astrophys. 382, 184 (2002) Crystalline silicate dust around evolved stars. I. The sample stars.

J/A+A/382/389/  clusters.dat table2.dat  (20Kb)

Kopylov A.I., Kopylova F.G. Astron. Astrophys. 382, 389 (2002) Search for streaming motion of galaxy clusters around the Giant Void

J/A+A/382/488/  table3.dat table4.dat  (11Kb)

Corsini E.M., Pizzella A., Bertola F. Astron. Astrophys. 382, 488 (2002) The orthogonal gaseous kinematical decoupling in the Sa spiral NGC 2855

J/A+A/382/513/  table1.dat table2.dat  (13Kb)

Reshetnikov V., Battaner E., Combes F., Jimenez-Vicente J. Astron. Astrophys. 382, 513 (2002) Statistics of galaxy warps in the HDF North and South.

J/A+A/382/899/  table1.dat table2.dat table3.dat tables.ps tables.tex  (309Kb)

Stassun K.G., van den Berg M., Mathieu R.D., Verbunt F. Astron. Astrophys. 382, 899 (2002) Photometric variability in the old open cluster M 67. II. General survey.

J/A+A/383/30/  ngc/* objects.dat table1.dat table9.dat table10.dat table11.dat  (178Kb)

Idiart T.P., Michard R., de Freitas Pacheco J.A. Astron. Astrophys. 383, 30 (2002) New UBVRI colour distributions in E-type galaxies. I. The data.

J/A+A/383/188/  paper.ps sp/* table1.dat table2.dat table3.dat tables1.ps  (14.3 Mb)

Munari U., Zwitter T. Astron. Astrophys. 383, 188 (2002) A multi-epoch spectrophotometric atlas of symbiotic stars.

J/A+A/383/197/  table2.dat  (12Kb)

Lyytinen J., Johansson P.,L.Jetsu, Esko E., Hackman T., Hall D.S., Henry G.W., Kontinen S., Koenoenen P., Maisala S., Palviainen A., Ryynaenen K. Astron. Astrophys. 383, 197 (2002) Time series analysis of V511 Lyrae photometry.

J/A+A/383/296/  table7.dat  (166Kb)

Peng Q., Vienne A., Shen K.X. Astron. Astrophys. 383, 296 (2002) Positional measuring procedure and CCD observations for Saturnian satellites.

J/A+A/383/398/  appendix.dat refs.dat table3.dat  (429Kb)

Paturel G., Theureau G., Fouque P., Terry J.N., Musella I., Ekholm T. Astron. Astrophys. 383, 398 (2002) Calibration of the distance scale from galactic Cepheids: I. Calibration based on the GFG sample.

J/A+A/383/580/  tableb.dat tables.tex tablev.dat  (29Kb)

Clarke D., McDavid D., Smith R.A., Henrichs H.F. Astron. Astrophys. 383, 580 (2002) On the polarimetric variability of bright O-type stars.

J/A+A/383/631/  cdn.dat  (828Kb)

Dutra C.M., Bica E. Astron. Astrophys. 383, 631 (2002) A catalogue of dust clouds in the Galaxy

J/A+A/383/719/  events.dat  (77Kb)

Arlot J.-E. Astron. Astrophys. 383, 719 (2002) Predictions of the mutual events of the Galilean Satellites of Jupiter occurring in 2002-2003

J/A+A/383/724/  innerjov.dat  (15Kb)

Kulyk I., Jockers K., Karpov N., Sergeev A. Astron. Astrophys. 383, 724 (2002) Astrometric CCD observations of the inner Jovian satellites in 1999-2000

J/A+A/383/823/  fcos.dat  (13Kb)

Mieske S., Hilker M., Infante L. Astron. Astrophys. 383, 823 (2002) Ultra compact objects in the Fornax cluster of galaxies: Globular clusters or dwarf galaxies?

J/A+A/383/1054/  table1.dat  (18Kb)

Rapaport M., Teixeira R., Le Campion J.F., Ducourant C., Camargo J.I.B., Benevides-Soares P. Astron. Astrophys. 383, 1054 (2002) Astrometry of Pluto and Saturn with the CCD meridian instruments of Bordeaux and Valinhos

J/A+A/383/1062/  table1.dat table2.dat  (16Kb)

Hu H., Wang R. Astron. Astrophys. 383, 1062 (2002) Optical positions of 55 radio stars from astrolabe observations of Yunnan Observatory

J/A+A/384/81/  lms.dat qso.dat vro.dat wd.dat  (55Kb)

Hatziminaoglou E., Groenewegen M.A.T., da Costa L., Arnouts S., Benoist C., Madejsky R., Mignani R., Olsen L.F., Rite C., Schirmer M., Slijkhuis R., Vandame B. Astron. Astrophys. 384, 81 (2002) ESO Imaging Survey. Deep Public Survey: Exploring the optical/infrared imaging data of CDF-S: Point

Sources.

J/A+A/384/145/  cc.dat  (5.8 Mb)

Belikov A.N., Kharchenko N.V., Piskunov A.E., Schilbach E., Scholz R.D., Yatsenko A.I. Astron. Astrophys. 384, 145 (2002) Study of the Per OB2 star forming complex. I. The compiled catalogue of kinematic and photometric data.

J/A+A/384/180 See I/276/

J/A+A/384/371/  table1.dat table2.dat table4.dat  (103Kb)

Simien, F., Prugniel, P. Astron. Astrophys. 384, 371 (2002) Kinematical data on early-type galaxies. VI.

J/A+A/384/383/  table4.dat table5.dat  (12Kb)

Iglesias-Paramo J., Boselli A., Cortese L., Vilchez J.M., Gavazzi G. Astron. Astrophys. 384, 383 (2002) A deep Hαsurvey of galaxies in the two nearby clusters Abell 1367 and Coma. The Hαluminosity functions.

J/A+A/384/473/  table3.dat table4.dat table5.dat table6.dat tables.ps tables.tex  (65Kb)

Lenorzer A., Vandenbussche B., Morris P., de Koter A., Geballe T., Waters L.B.F.M., Hony S., Kaper L. Astron. Astrophys. 384, 473 (2002) An atlas of 2.4 to 4.1 mum ISO/SWS spectra of early-type stars

J/A+A/384/491/ figure1.ps figure2.ps figure3.ps figure4.ps figure5.ps figure6.ps figure7.ps figure8.ps notes.dat table1.dat table3.dat table4.dat table5.dat (398Kb) Cutispoto G., Pastori L., Pasquini L., de Medeiros J.R., Tagliaferri G., Andersen J. Astron. Astrophys. 384, 491 (2002) II. Li abundances, vsini and X-ray luminosities relationships. Cutispoto G., Tagliaferri G., de Medeiros J.R., Pastori L., Pasquini L., Andersen J. Astron. Astrophys. 397, 987 (2003) Fast-rotating nearby solar-type stars. I. Spectral classification, vsini, Li abundances and X-ray luminosities

J/A+A/384/504/  table2.dat  (11Kb)

Evans A., Yudin R.V., Naylor T., Ringwald F.A., Koch Miramond L. Astron. Astrophys. 384, 504 (2002) Broadband polarimetry of novae in outburst.

J/A+A/384/650/  table.dat  (16Kb)

Penna J.L., Jilinski E.G., Andrei A.H., Puliaev S.P., Reis Neto Astron. Astrophys. 384, 650 (2002) CCD astrolabe solar diameter measurements in 1999-2000

J/A+A/384/879/  table1.dat table2.dat table4.dat  (18Kb)

Barbieri M., Gratton R.G. Astron. Astrophys. 384, 879 (2002) Galactic orbits of stars with planets

J/A+A/384/890/  table3.dat table3.tex  (51Kb)

Berghoefer T.W., Christian D.J. Astron. Astrophys. 384, 890 (2002) Optical properties of X-ray selected stars in NGC 2244 in the Rosette Nebula.

J/A+A/384/925/  catalog.dat  (144Kb)

Kharchenko N., Kilpio E., Malkov O., Schilbach E. Astron. Astrophys. 384, 925 (2002) Mira kinematics in the post-Hipparcos era.

J/A+A/384/1012/ table3.dat table4.dat table5.dat table6.dat table8.dat table9.dat table10.dat table11.dat table12.dat table13.dat (34Kb) Paardekooper S.J., Veen P.M., van Genderen A., Van der Hucht K.A. Astron. Astrophys. 384, 1012 (2002) On the variability of the visual binary WR 86. WC 7 with a beta-Cephei companion

J/A+A/384/1030/  table4.dat  (8Kb)

Sterzik M.F., Tokovinin A.A. Astron. Astrophys. 384, 1030 (2002) Relative orientation of orbits in triple stars.

J/A+A/384/1050/  table1.dat  (10Kb)

Berdyugin A., Teerikorpi P. Astron. Astrophys. 384, 1050 (2002) Interstellar polarization at high galactic latitudes from distant stars. VI. Extended polarization map and

connection with the local spiral structure.

J/A+A/385/39/  table3.dat table4.dat  (21Kb)

Venturi T., Bardelli S., Zagaria M., Prandoni I., Morganti R. Astron. Astrophys. 385, 39 (2002) Radio emission from the A3571 cluster complex: The final stage of a cluster merger?

J/A+A/385/55/  table1.dat  (36Kb)

Hagen-Thorn V.A., Larionova E.G., Jorstad S.G., Bjornsson C.-I., Larionov V.M. Astron. Astrophys. 385, 55 (2002) Analysis of the long-term polarization behaviour of BL Lac.

J/A+A/385/67/  tablea1.dat tablea2.dat tablea3.dat  (59Kb)

Feltzing S., Johnson R.A. Astron. Astrophys. 385, 67 (2002) A new, cleaner colour-magnitude diagram for the metal-rich globular cluster NGC 6528. Velocity dispersion in the Bulge, age and proper motion of NGC 6528.

J/A+A/385/87/  notes.dat table1.dat table1.ps table1.tex table2.dat table2.ps table2.tex  (127Kb)

Balega I.I., Balega Y.Y., Hofmann K.-H., Maksimov A.F., Pluzhnik E.A., Schertl D., Shkhagosheva Z.U., Weigelt G. Astron. Astrophys. 385, 87 (2002) Speckle interferometry of nearby multiple stars.

J/A+A/385/377/  table1.dat  (9Kb)

Quilligan F., McBreen B., Hanlon L., McBreen S., Hurley K.J., Watson D. Astron. Astrophys. 385, 377 (2002) Temporal properties of gamma ray bursts as signatures of jets from the central engine.

J/A+A/385/471/  ngc7036.dat ngc7772.dat  (45Kb)

Carraro G. Astron. Astrophys. 385, 471 (2002) Photometry of dissolving star cluster candidates. The cases of NGC 7036 and NGC 7772.

J/A+A/385/488/  rv.dat table3.dat vmag.dat  (38Kb)

Carrier F., Burki G., Burnet M. Astron. Astrophys. 385, 488 (2002) Search for duplicity in periodic variable Be stars

J/A+A/385/716/  table4.dat table4.tex table5.dat table5.tex  (694Kb)

Tachiev G.I., Froese Fischer C. Astron. Astrophys. 385, 716 (2002) Breit-Pauli energy levels and transition rates for nitrogen-like and oxygen-like sequences.

J/A+A/385/793/  table3.dat table4.dat table5.dat  (13Kb)

Antonello E., Fugazza D., Mantegazza L. Astron. Astrophys. 385, 793 (2002) Search for second overtone mode Cepheids in Magellanic Clouds. II. Characteristics of second overtone mode pulsations.

J/A+A/385/816/  galaxies.dat procurve.dat rawcurve.dat  (112Kb)

de Blok W.J.G., Bosma A. Astron. Astrophys. 385, 816 (2002) High-resolution rotation curves of low surface brightness galaxies

J/A+A/385/867/  table4.dat table5.dat table6.dat table7.dat table8.dat  (61Kb)

Paunzen E., Maitzen H.M. Astron. Astrophys. 385, 867 (2002) CCD photometric search for peculiar stars in open clusters. III. NGC 2439, NGC 3960, NGC 6134, NGC 6192 and NGC 6451.

J/A+A/385/884/  notes.dat table1.dat table2.dat table3.dat table4.dat  (31Kb)

Fujii T., Nakada Y., Parthasarathy M. Astron. Astrophys. 385, 884 (2002) BVRIJHK photometry of post-AGB candidates.

J/A+A/385/986/  table2.dat  (15Kb)

Fremat Y., Zorec J., Hubert A.-M., Cidale L.S., Rohrmann R.D., Desert J.-M., Ferlet R., Vidal-Madjar A. Astron. Astrophys. 385, 986 (2002) FUSE observations towards the pole-on Be star HR 5223.

J/A+A/385/1014/  table2.dat  (20Kb)

Chrysostomou, A., Aitken, D.K., Jenness, T., Davis, C.J., Hough, J.H., Curran, R., Tamura, M. Astron. Astrophys. 385, 1014 (2002) The magnetic field structure in W51A

J/A+A/386/97/  notes.dat refs.dat tablea1.dat  (154Kb)

Jackson C.A., Wall J.V., Shaver P.A., Kellermann K.I., Hook I.M., Hawkins M.R.S. Astron. Astrophys. 386, 97 (2002) The Parkes quarter-Jansky flat-spectrum sample. I. Sample selection and source identifications.

J/A+A/386/140/  table1.dat  (11Kb)

van Driel W., Combes F., Arnaboldi M., Sparke S. Astron. Astrophys. 386, 140 (2002) A neutral hydrogen survey of polar ring galaxies. IV. Parkes observations

J/A+A/386/187/  table1.dat table2.dat tablea1.dat  (38Kb)

Meibom S., Andersen J., Nordstroem B. Astron. Astrophys. 386, 187 (2002) Critical tests of stellar evolution in open clusters. III: stellar population and dynamical evolution of IC 4651.

J/A+A/386/237/  table2.dat  (31Kb)

Munari U., Jurdana-Sepic R. Astron. Astrophys. 386, 237 (2002) Symbiotic stars on Asiago archive plates. II.

J/A+A/386/492/  charm.dat refs.dat  (1.3 Mb)

Richichi A., Percheron I. Astron. Astrophys. 386, 492 (2002) CHARM: A Catalog of High Angular Resolution Measurements.

J/A+A/386/709 See IV/25/

J/A+A/386/1001/  fe.dat h.dat sol.dat  (10.6 Mb)

Gaensicke, B.T., Braje, T.M., Romani, R.W. Astron. Astrophys. 386, 1001 (2002) Thermal emission from low-field neutron stars

J/A+A/387/1/  catalog.dat  (127Kb)

Vauglin I., Rousseau J., Paturel G., Borsenberger J., Epchtein N., Fouque P., Kimeswenger S., Le Bertre T., Mamon G.A. Astron. Astrophys. 387, 1 (2002) Serendipitous detection of galaxies behind the Milky Way from the DENIS survey.

J/A+A/387/26/  table6.dat table7.dat  (59Kb)

Fasano G., Bettoni D., D’Onofrio M., Kjfrgaard P., Moles M. Astron. Astrophys. 387, 26 (2002) The scaling relations of early-type galaxies in clusters. I. Surface photometry in seven nearby clusters

J/A+A/387/129/  table1.dat table2.dat table3.dat  (28Kb)

Wang X.-H., Chen P.-S. Astron. Astrophys. 387, 129 (2002) Near-infrared observations of candidate extrinsic S stars.

J/A+A/387/240/  v899her.dat  (22Kb)

Ozdemir S., Demircan O., Erdem A., Cicek C., Bulut I., Soydugan F., Soydugan E. Astron. Astrophys. 387, 240 (2002) A photometric study of the recently discovered eclipsing binary V899 Herculis

J/A+A/387/479/  ngc133.dat ngc1348.dat  (43Kb)

Carraro G. Astron. Astrophys. 387, 479 (2002) A photometric study of the two poorly know northern open clusters NGC 133 and NGC 1348

J/A+A/387/580/  table3.dat table5.dat table6.dat table7.dat table567.ps table567.tex  (187Kb)

Harmanec P., Bozic H., Percy J., Yang S., Ruzdjak D., Sudar D., Wolf M., Illiev L., Huang L., Buil C., Eenens P. Astron. Astrophys. 387, 580 (2002) Properties and nature of Be stars 21. The long-term and the orbital variations of V832 Cyg = 59 Cyg

J/A+A/387/850/  table1.dat  (47Kb)

Imbert M. Astron. Astrophys. 387, 850 (2002) Vitesses radiales photoelectriques de binaires a eclipses. VI. Orbites spectroscopiques et elements physiques de 12 etoiles doubles. Photoelectric radial velocities of eclipsing binaries. VI. Orbital and physical elements of 12 double stars.

J/A+A/387/1135/ table4a.dat table4b.dat table4c.dat table4d.dat table5.dat table6.dat table7.dat table8.dat (113Kb) Kisielius R., Storey P.J. Astron. Astrophys. 387, 1135 (2002) Effective recombination coefficients for NII lines at nebular temperatures and densities.

J/A+A/387/1146/  tables.dat  (13Kb)

Charro E., Curiel Z., Martin I. Astron. Astrophys. 387, 1146 (2002) Atomic data for M1 and E2 emission lines in the potassium isoelectronic sequence

J/A+A/388/1/  tablea1.dat tablea2.dat tablea3.dat tablea4.dat tablea5.dat  (15Kb)

Hansen L., Olsen L.F., Jorgensen H.E. Astron. Astrophys. 388, 1 (2002) Multi-object spectroscopy of low redshift EIS clusters. I.

J/A+A/388/29/  table3.dat  (14Kb)

Parodi B.R., Barazza F.D., Binggeli B. Astron. Astrophys. 388, 29 (2002) Structure and stellar content of dwarf galaxies. VII. B and R photometry of 25 southern field dwarfs and a disk parameter analysis of the complete sample of nearby irregulars.

J/A+A/388/100/  notes.dat refs.dat table1.dat table3.dat  (55Kb)

Kahabka P. Astron. Astrophys. 388, 100 (2002) ROSAT X-ray sources in the field of the LMC. II. Statistics of background AGN and X-ray binaries

J/A+A/388/149/  table5.dat  (25Kb)

Smriglio F., Dasgupta A.K., Boyle R.P. Astron. Astrophys. 388, 149 (2002) CCD photometry in the Vilnius system in the vicinity of the globular cluster M 56.

J/A+A/388/168/  notes.dat pm/* table1.dat  (746Kb)

Dias W.S., Lepine J.R.D., Alessi B.S. Astron. Astrophys. 388, 168 (2002) Proper motions of open clusters based on the TYCHO2 Catalogue. II. Clusters farther than 1 kpc.

J/A+A/388/179/  table2.dat table3.dat  (61Kb)

Piatti A.E., Claria J.J. Astron. Astrophys. 388, 179 (2002) Two highly reddened young open clusters located beyond the Sagittarius arm

J/A+A/388/298/  table1.dat  (19Kb)

Frasca A., Cakirli O., Catalano S., Ibanoglu C., Marilli E., Evren S., Tas G. Astron. Astrophys. 388, 298 (2002) Hαspectroscopy and BV photometry of RT Lacertae

J/A+A/388/632/  hd108147.dat hd168746.dat  (17Kb)

Pepe F., Mayor M., Galland F., Naef D., Queloz D., Santos N.C., Udry S., Burnet M. Astron. Astrophys. 388, 632 (2002) The CORALIE survey for southern extra-solar planets. VII. Two short-period Saturnian companions to HD 108147 and HD 168746.

J/A+A/388/704/  table2.dat table3.dat table4.dat table5.dat  (20Kb)

Vujnovic V., Blagoev K., Fuerboeck C., Neger T., Jaeger H. Astron. Astrophys. 388, 704 (2002) Absolute transition probabilities of Al I and Al II spectral lines and intensity ratios within multiplets.

J/A+A/389/149/  lcurve/* lightcur.dat longper.dat shortper.dat  (1.6 Mb)

Derue F., Marquette J-B., Lupone S., Afonso C., Alard C., Albert J-N., Amadon A., Andersen J., Ansari R., Aubourg E., Bareyre P., Bauer F., Beaulieu J-P., Blanc G., Bouquet A., Char S., Charlot X., Couchot F., Coutures C., Ferlet R., Fouque P., Glicenstein J-F., Goldman B., Gould A., Graff D., Gros M., Haissinski J., Hamilton J-C., Hardin D., de Kat J., Kim A., Lasserre T., Le Guillou L., Lesquoy E., Loup C., Magneville C., Mansoux B., Maurice E., Milsztajn A., Moniez M., Palanque-Delabrouille N., Perdereau O., Prevot L., Regnault N., Rich J., Spiro M., Vidal-Madjar A., Vigroux L., Zylberajch S. The EROS collaboration Astron. Astrophys. 389, 149 (2002) Observations of periodic variable stars towards the Galactic spiral arms by EROS II.

J/A+A/389/439/  table3.dat table3.ps table3.tex  (59Kb)

Nedialkov P., Orio M., Birkle K, Conselice C., Della Valle M., Greiner J., Magnier E., Tikhonov N.A. Astron. Astrophys. 389, 439 (2002) First optical identification of a supersoft X-ray source in M31.

J/A+A/389/641/  table1.dat  (17Kb)

Hainaut O.R., Delsanti A.C. Astron. Astrophys. 389, 641 (2002) Colors of Minor Bodies in the Outer Solar System. A statistical analysis

J/A+A/389/716/  en/* fjj/* table3a.dat table3b.dat table7.dat target.dat  (2.6 Mb)

Nahar S. Astron. Astrophys. 389, 716 (2002) Relativistic fine structure oscillator strengths for Li-like ions: CIV - Si XII, S XIV, Ar XVI, Ca XVIII, Ti XX, Cr XXII, and Ni XXVI

J/A+A/389/787/  ac118.dat  (204Kb)

Busarello G., Merluzzi P., La Barbera F., Massarotti M., Capaccioli M. Astron. Astrophys. 389, 787 (2002) Optical (VRI) Photometry in the Field of the Galaxy Cluster AC118 at z=0.31.

J/A+A/389/795/  table1.dat  (9Kb)

Bizyaev D., Mitronova S. Astron. Astrophys. 389, 795 (2002) Photometric parameters of edge-on galaxies from 2MASS observations.

J/A+A/389/845/  fig4-1.ps fig4-2.ps fig4-3.ps fig4-4.ps fig4-5.ps fig4-6.ps fig4-7.ps table3.dat  (3.0 Mb)

Kong X., Cheng F.Z. Astron. Astrophys. 389, 845 (2002) Spectroscopic study of blue compact galaxies. I. The spectra.

J/A+A/389/931/  table1.dat table2.dat  (10Kb)

Tubbesing S., Kaufer A., Stahl O., Wolf B., Schmid H.M., Korn A.J., Maintz M., Rivinius T., Szeifert T., Arentoft T., Sterken C. Astron. Astrophys. 389, 931 (2002) The eclipsing hypergiant R 81 (B2.5Ia-O) in the Large Magellanic Cloud. System properties from spectroscopic and photometric monitoring.

J/A+A/390/39/  table1.dat table2.dat  (20Kb)

Aharonian F., Akhperjanian A., Barrio J.A., Bernloehr K., Boerst H., Bojahr H., Bolz O., Contreras J.L., Cortina J., Denninghoff S., Fonseca V., Gebauer H.J., Gonzalez J., Goetting N., Heinzelmann G., Hermann G., Heusler A., Hofmann W., Horns D., Jung I., Kankanyan R., Kestel M., Kettler J., Kohnle A., Konopelko A., Kornmeyer H., Kranich D., Krawczynski H., Lampeitl H., Lopez M., Lorenz E., Lucarelli F., Magnussen N., Mang O., Meyer H., Mirzoyan R., Moralejo A., Ona E., Padilla L., Panter M., Plaga R., Plyasheshnikov A., Prahl J., Puehlhofer G., Rauterberg G., Roehring A., Rhode W., Rowell G., Sahakian V., Samorski M., Schilling M., Schmele D., Schroeder F., Sevilla I., Siems M., Stamm W., Tluczykont M., Voelk H.J., Wiedner C.A., Wittek W. Astron. Astrophys. 390, 39 (2002) Search for point sources of gamma radiation above 15 TeV with the HEGRA AIROBICC array.

J/A+A/390/173/  table4.dat  (8Kb)

Mennessier M.O., Colome J. Astron. Astrophys. 390, 173 (2002) Identification of RR Lyrae stars in the Tycho-2 catalogue.

J/A+A/390/179/  table1.dat  (11Kb)

Korhonen H., Berdyugina S.V., Tuominen I. Astron. Astrophys. 390, 179 (2002) Study of FK Comae Berenices. IV. Active longitudes and the ”flip-flop” phenomenon

J/A+A/390/235/  stars.dat tablea1.dat  (12Kb)

Nissen P.E., Primas F., Asplund M. Lambert D.L. Astron. Astrophys. 390, 235 (2002) O/Fe in metal-poor main sequence and subgiant stars.

J/A+A/390/267/  hd141937.dat hd162020.dat hd168443.dat hd202206.dat  (17Kb)

Udry S., Mayor M., Naef D., Pepe F., Queloz D., Santos N.C., Burnet M. Astron. Astrophys. 390, 267 (2002) The CORALIE survey for southern extra-solar planets. VIII. The very low-mass companions of HD 141937, HD 162020, HD 168443 and HD 202206 : Brown dwarfs or ”superplanets”?

J/A+A/390/491/  table1.dat  (25Kb)

Reyle C., Robin A.C., Scholz R.-D., Irwin M. Astron. Astrophys. 390, 491 (2002) New nearby stars selected in a high proper motion survey by DENIS photometry.

J/A+A/390/501/  table1.dat table3.dat  (58Kb)

Groenewegen M.A.T., Sevenster M., Spoon H.W.W., Perez I. Astron. Astrophys. 390, 501 (2002) Millimetre Observations of Infrared Carbon Stars I: The Data

J/A+A/390/511/  table1.dat  (29Kb)

Groenewegen M.A.T., Sevenster M., Spoon H.W.W., Perez I. Astron. Astrophys. 390, 511 (2002) Millimetre observations of infrared carbon stars. II. Mass loss rates and expansion velocities.

J/A+A/390/585/  rcygall.dat rcygavg.dat signal.dat  (523Kb)

Kiss L.L., Szatmary K. Astron. Astrophys. 390, 585 (2002) Period-doubling events in the light curve of R Cygni: Evidence for chaotic behaviour.

J/A+A/390/707/  catalog.dat  (15Kb)

Temmer, M., Veronig, A., Hanslmeier, A. Astron. Astrophys. 390, 707 (2002) Hemispheric Sunspot Numbers Rn and Rs: Catalogue and N-S asymmetry analysis

J/A+A/390/717/ v89f.dat v90f.dat v91f.dat v92f.dat v93f.dat v94f.dat v95f.dat v96f.dat v97f.dat v98f.dat v99f.dat (244.1 Mb) Salabert D., Fossat E., Gelly B., Tomczyk S., Palle P.L., Jimenez-Reyes S.J., Cacciani A., Corbard T., Ehgamberdiev S., Grec G., Hoeksema J.T., Kholikov S., Lazrek M., Schmider F.X. Astron. Astrophys. 390, 717 (2002) IRIS++ database: Merging of IRIS + Mark-1 + LOWL

J/A+A/390/863/ refs2.dat tablea1.dat tablea2.dat tablea3.dat tablea4.dat tablea5.dat tablea6.dat (70Kb) Swaters R.A., Balcells M. Astron. Astrophys. 390, 863 (2002) The Westerbork HI survey of spiral and irregular galaxies. II. R-band surface photometry of late-type dwarf galaxies.

J/A+A/390/931/  pismis23.dat stephen2.dat  (365Kb)

Ortolani S., Bica E., Barbuy B., Momany Y. Astron. Astrophys. 390, 931 (2002) The very reddened open clusters Pismis 23 (Lynga 10) and Stephenson 2.

J/A+A/390/961/  catalog.dat  (22Kb)

Andreuzzi G., Richer H.B., Limongi M., Bolte M. Astron. Astrophys. 390, 961 (2002) The white dwarf cooling sequence in the old open cluster NGC 188.

J/A+A/390/967/  table2.dat  (42Kb)

Bergeat J., Knapik A., Rutily B. Astron. Astrophys. 390, 967 (2002) Carbon-rich giants in the HR diagram and their luminosity function

J/A+A/390/1023/  table2.dat table3.dat table4.dat table5.dat tablea.dat  (34Kb)

Adelman S.J., Meadows S.A. Astron. Astrophys. 390, 1023 (2002) uvby Photometry of the CP stars HR 194, HD 32966, HD 171782, and HR 7911

J/A+A/390/1033/  table1.dat  (33Kb)

Vandenbussche B., Beintema D., de Graauw T., Decin L., Feuchtgruber H., Heras A., Kester D., Lahuis F., Lenorzer A., Lorente R., Salama A., Waelkens C., Waters L., Wieprecht E. Astron. Astrophys. 390, 1033 (2002) The ISO-SWS post-helium atlas of near-infrared stellar spectra

J/A+A/391/L11/  lyalpha.dat lybeta.dat lyman.dat stis.dat uves.dat  (24Kb)

Janknecht E., Baade R., Reimers D. Astron. Astrophys. 391, L11 (2002) A high-resolution study of the evolution of the Lyman alpha forest in the redshift interval 0.9 z 1.7.

J/A+A/391/35/  table1.dat table1.tex table2.dat  (98Kb)

Focardi P., Kelm B.

Astron. Astrophys. 391, 35 (2002) Compact Groups in the UZC galaxy sample.

J/A+A/391/159/  table1.dat  (112Kb)

De Heij V., Braun R., Burton W.B. Astron. Astrophys. 391, 159 (2002) An automated search for compact high-velocity clouds in the Leiden/Dwingeloo Survey

J/A+A/391/179/  table.dat  (14Kb)

Ortolani S., Carraro G., Covino S., Bica E., Barbuy B. Astron. Astrophys. 391, 179 (2002) A photometric study of the young open cluster NGC 1220.

J/A+A/391/481/  table2.dat  (7Kb)

Burud I., Hjorth J., Courbin F., Cohen J.G., Magain P., Jaunsen A.O., Kaas A.A., Faure C., Letawe G. Astron. Astrophys. 391, 481 (2002) Time delay and lens redshift for the doubly imaged BAL quasar SBS1520+530

J/A+A/391/519/  refs.dat table1.dat table2.dat  (22Kb)

Castro-Rodriguez N., Lopez-Corredoira M., Sanchez-Saavedra M.L., Battaner E. Astron. Astrophys. 391, 519 (2002) Warps and correlations with intrinsic parameters of galaxies in the visible and radio.

J/A+A/391/547/  clusters.dat refs.dat  (179Kb)

Dieball A., Mueller H., Grebel E.K. Astron. Astrophys. 391, 547 (2002) A statistical study of binary and multiple clusters in the LMC

J/A+A/391/641/  mg2.dat table3a.dat table3b.dat table6.dat  (24Kb)

Ekmekci F., Karakus O. Astron. Astrophys. 391, 641 (2002) IUE-newsips spectra of σGeminorum

J/A+A/391/767/  phoebe.dat  (13Kb)

Fienga A., Arlot J.-E., Baron N., Bec-Borsenberger A., Crochot A., Emelyanov N., Thuillot W. Astron. Astrophys. 391, 767 (2002) CCD observations of Phoebe, 9th satellite of Saturn, made in 1998 and 1999 at Observatoire de Haute Provence.

J/A+A/391/775/  table2.dat tablea1.dat  (36Kb)

Shen K.X., Qiao R.C., Harper D., Hadjifotinou K.G., Liu J.R. Astron. Astrophys. 391, 775 (2002) Astrometry of five Uranian satellites in 1995 and 1997

J/A+A/391/809/  table3.dat table4.dat table5.dat table6.dat table7.dat table8.dat  (86Kb)

Cedres B., Cepa J. Astron. Astrophys. 391, 809 (2002) Distributions, equivalent widths and metallicities of the HII regions in the spiral galaxies NGC 5457 and NGC 4395.

J/A+A/391/857/  table2.dat  (9Kb)

Homeier N., Gallagher III J.S., Pasquali A. Astron. Astrophys. 391, 857 (2002) The star cluster system of the NGC 7673 starburst.

J/A+A/391/887/  table1.dat table2.dat  (34Kb)

Kraan-Korteweg R.C., Henning P.A., Schroeder A.C. Astron. Astrophys. 391, 887 (2002) Parkes HI observations of galaxies. behind the southern Milky Way. I. The Hydra/Antlia region (l= 266^{∘}to 296^{∘}).

J/A+A/391/945/  db/* table2.dat  (36.3 Mb)

Piotto G., King I.R., Djorgovski S.G., Sosin C., Zoccali M., Saviane I., de Angeli F., Riello M., Recio-Blanco A., Rich R.M., Meylan G., Renzini A. Astron. Astrophys. 391, 945 (2002) HST color-magnitude diagrams of 74 galactic globular clusters in the HST F439W and F555W bands.

J/A+A/391/967/  table2.dat table3.dat table4.dat  (17Kb)

Sjouwerman L.O., Lindqvist M., van Langevelde H.J., Diamond P.J. Astron. Astrophys. 391, 967 (2002) H2O and SiO maser emission in Galactic center OH/IR stars.

J/A+A/391/1039/  table2.dat table3.dat  (67Kb)

Kinman T., Castelli F. Astron. Astrophys. 391, 1039 (2002) The determination of Teff for metal-poor A-type stars using V and 2MASS J, H and K magnitudes

J/A+A/392/1/ saecoeff.dat sahst.dat sajohn.dat sakatt.dat sakcoeff.dat samag.dat saplot1.dat saplot2.dat saplot3.dat saplot4.dat saspect.dat scecoeff.dat schst.dat scjohn.dat sckatt.dat sckcoeff.dat scmag.dat scplot1.dat scplot2.dat scplot3.dat scplot4.dat scspect.dat (33.6 Mb) Schulz J., Fritze-v.Alvensleben U., Moeller C.S. Astron. Astrophys. 392, 1 (2002) Spectral + photometric evolution of Simple Stellar Populations at various metallicities

J/A+A/392/215/  table1.dat  (7Kb)

Santos N.C., Mayor M., Naef D., Pepe F., Queloz D., Udry S., Burnet M., Clausen J.V., Helt B.E., Olsen E.H., Pritchard J.D. Astron. Astrophys. 392, 215 (2002) The CORALIE survey for southern extrasolar planets. IX. A 1.3-day period brown dwarf disguised as a planet

J/A+A/392/239/  notes.dat table3.dat  (38Kb)

Stanke T., McCaughrean M.J., Zinnecker H. Astron. Astrophys. 392, 239 (2002) An unbiased H_{2} survey for protostellar jets in Orion A. II. The infrared survey data.

J/A+A/392/395/  0055k.dat cdfsk.dat  (32Kb)

Cimatti A., Mignoli M., Daddi E., Pozzetti L., Fontana A., Saracco P., Poli F., Renzini A., Zamorani G., Broadhurst T., Cristiani S., D’Odorico S., Giallongo E., Gilmozzi R., Menci N. Astron. Astrophys. 392, 395 (2002) The K20 survey. III. Photometric and spectroscopic properties of the sample.

J/A+A/392/491/  objects.dat objects.tex tablea1.dat tablea1.ps tablea1.tex  (114Kb)

Andrievsky S.M., Kovtyukh V.V., Luck R.E., Lepine J.R.D., Maciel W.J., Beletsky Yu.V. Astron. Astrophys. 392, 491 (2002) Using Cepheids to determine the galactic abundance gradient. III. First results for the outer disc

J/A+A/392/741/  table3.dat table4.dat table5.dat table6.dat table7.dat table8.dat  (236Kb)

Groenewegen M.A.T., Girardi L., Hatziminaoglou E., Benoist C., Olsen L.F., da Costa L., Arnouts S., Madejsky R., Mignani R.P., Rite C., Sikkema G., Slijkhuis R. Astron. Astrophys. 392, 757 (2002) ESO Imaging Survey. The Stellar Catalogue in the Chandra Deep Field South

J/A+A/392/795/  table3.dat table4.dat table11.dat table12.dat  (22Kb)

Valtchanov I., Murphy T., Pierre M., Hunstead R., Lemonon L. Astron. Astrophys. 392, 795 (2002) Abell 1451 and 1RXS J131423.6-251521: A multi-wavelength study of two dynamically perturbed clusters of galaxies.

J/A+A/392/851/  table3.dat table4.dat  (24Kb)

Meusinger H., Scholz R.-D., Irwin M., Lehmann H. Astron. Astrophys. 392, 851 (2002) QSOs from the variability and proper motion survey in the M3 field.

J/A+A/392/971/  table3.dat  (83Kb)

Felli M., Testi L., Schuller F., Omont A. Astron. Astrophys. 392, 971 (2002) Young massive stars in the ISOGAL survey. II. The catalogue of bright YSO candidates.

J/A+A/392/1129/  tablea1.dat tablea2.dat tablea3.dat  (27Kb)

Ng Y.K., Brogt E., Chiosi C., Bertelli G. Astron. Astrophys. 392, 1129 (2002) Automatic observation rendering (AMORE). I. On a synthetic stellar population’s colour-magnitude diagram.

J/A+A/393/57/  table2.dat  (23Kb)

Terry J.N., Paturel G., Ekholm T. Astron. Astrophys. 393, 57 (2002) Local velocity field from sosie galaxies I: The Peebles’model.

J/A+A/393/115/  table2.dat table3.dat  (38Kb)

Messineo M., Habing H.J., Sjouwerman L.O., Omont A., Menten K.M. Astron. Astrophys. 393, 115 (2002) 86 GHz SiO maser survey of late-type stars in the Inner Galaxy. I. Observational data

J/A+A/393/167/ bins.dat c b5 4.dat c b5 8.dat c b5 12.dat c b5 16.dat c b5 20.dat c bin1.dat c bin2.dat c bin3.dat c bin4.dat c bin5.dat c co101.dat c co110.dat c co140.dat o b9 2.dat o b9 4.dat o b9 6.dat o b9 8.dat o b9 10.dat o b9 12.dat o b9 14.dat o b9 16.dat o b9 18.dat o b9 20.dat o b9 24.dat o b9 30.dat o b9 40.dat o bin1.dat o bin2.dat o bin3.dat o bin4.dat o bin5.dat o bin6.dat o bin7.dat o bin8.dat o bin9.dat table12.dat (2.9 Mb) Lancon A., Mouhcine M. Astron. Astrophys. 393, 167 (2002) The modelling of intermediate-age stellar populations. II. Average spectra for upper AGB stars, and their use.

J/A+A/393/183/  catalog.dat  (129Kb)

Borde P., Coude du Foresto V., Chagnon G., Perrin G. Astron. Astrophys. 393, 183 (2002) A catalogue of calibrator stars for long baseline stellar interferometry

J/A+A/393/225/  table3.dat table4.dat table5.dat table6.dat table7.dat table8.dat  (22Kb)

Messina S., Guinan E.F. Astron. Astrophys. 393, 225 (2002) Magnetic activity of six young solar analogues. I. Starspot cycles from long-term photometry.

J/A+A/393/389/  table5.dat  (41Kb)

Marquez I., Masegosa J., Moles M., Varela J., Bettoni D., Galletta G. Astron. Astrophys. 393, 389 (2002) Rotation curves and metallicity gradients from H II regions in spiral galaxies.

J/A+A/393/425/  table2.dat  (18Kb)

Mainieri V., Bergeron J., Hasinger G., Lehmann I., Rosati P., Schmidt M., Szokoly G., Della Ceca R. Astron. Astrophys. 393, 425 (2002) XMM-Newton observation of the Lockman Hole. II. Spectral analysis.

J/A+A/393/439/  catalog.dat  (4.0 Mb)

Durret F., Adami C., Lobo C. Astron. Astrophys. 393, 439 (2002) Environmental effects on the galaxy luminosity function in the cluster of galaxies Abell 496.

J/A+A/393/897/  table5.dat table9.dat  (83Kb)

Royer F., Grenier S., Baylac M.-O., Gomez A.E., Zorec J. Astron. Astrophys. 393, 897 (2002) Rotational velocities of A-type stars. II. Measurement of vsini in the northern hemisphere.

J/A+A/393/965/  table1.dat tables.dat  (26Kb)

de Cat P., Aerts C. Astron. Astrophys. 393, 965 (2002) A study of bright southern slowly pulsating B stars. II. The intrinsic frequencies.

J/A+A/393/1053/  comete.tar index.dat spectra/* tablesa.dat tablesb.dat tablesc.dat  (27.5 Mb)

Crovisier J., Colom P., Gerard E., Bockelee-Morvan D., Bourgois G. Astron. Astrophys. 393, 1053 (2002) Observations at Nancay of the OH 18-cm lines in comets. The data base. Observations made from 1982 to 1999.

J/A+A/394/59 See VIII/69A/

J/A+A/394/97/  b-v.dat b-y.dat c1mag.dat deltav.dat m1mag.dat u-b.dat vmag.dat  (330Kb)

Kiss L.L., Derekas A., Alfaro E.J., Biro I.B., Csak B., Garrido R., Szatmary K., Thomson J.R. Astron. Astrophys. 394, 97 (2002) The multimode pulsation of the delta Scuti star V784 Cassiopeae.

J/A+A/394/151/  stars.dat table1.dat table2.dat table3.dat  (145Kb)

Carrier F., North P., Udry S., Babel J. Astron. Astrophys. 394, 151 (2002) Multiplicity among chemically peculiar stars. II. Cool magnetic Ap stars.

J/A+A/394/187/  table3a.dat table3b.dat table4.dat  (58Kb)

Caliskan H., Adelman S.J., Cay M.T., Gulliver A.F., Tektunali H.G., Teker A. Astron. Astrophys. 394, 187 (2002) Elemental abundance analyses with DAO spectrograms. XXVI. The superficially normal stars γSerpentic (F6 V) and 101 Herculis (A7 V).

J/A+A/394/225/  table1.dat  (20Kb)

Zavagno A., Deharveng L., Nadeau D., Caplan J. Astron. Astrophys. 394, 225 (2002) Near-infrared imaging of RAFGL 7009S.

J/A+A/394/395/  catalog.dat  (29Kb)

Dietrich, J.P., Clowe, D.I., Soucail, G. Astron. Astrophys. 394, 395 (2002) Spectroscopy of the neighboring massive clusters Abell 222 and Abell 223

J/A+A/394/479/  table6.dat table7.dat  (63Kb)

Capilla, G., Fabregat J. Astron. Astrophys. 394, 479 (2002) CCD uvbyβphotometry of young open clusters. I. The double cluster h and χPersei.

J/A+A/394/505/  table2.dat  (39Kb)

Berdyugina S.V., Pelt J., Tuominen I. Astron. Astrophys. 394, 505 (2002) Magnetic activity in the young solar analog LQ Hydrae. I. Active longitudes and cycles.

J/A+A/394/617/  table1.dat table2.dat  (54Kb)

Yudin B.F., Fernie J.D., Ikhsanov N.R., Shenavrin V.I., Weigelt G. Astron. Astrophys. 394, 617 (2002) UBVJHKLM photometry and modeling of R Corona Borealis

J/A+A/394/753/ (3Kb) Storey P.J., Zeippen C.J., Le Dourneuf M. Astron. Astrophys. 394, 753 (2002) Atomic data from the IRON Project. LI. Electron impact excitation of Fe IX.

J/A+A/394/927/  table2.dat table3.dat  (120Kb)

Ibukiyama A., Arimoto N. Astron. Astrophys. 394, 927 (2002) HIPPARCOS age-metallicity relation of the solar neighbourhood disc stars.

J/A+A/394/943/  v567oph.dat  (102Kb)

Kiss L.L., Derekas A., Meszaros Sz., Szekely P. Astron. Astrophys. 394, 943 (2002) A photometric monitoring of bright high-amplitude delta Scuti stars. I. The double-mode pulsation of V567 Ophiuchi

J/A+A/394/975/  table1.dat table1.ps table1.tex  (68Kb)

Desmurs J.-F., Baudry A., Sivagnanam P., Henkel C. Astron. Astrophys. 394, 975 (2002) A high-sensitivity OH 5-cm line survey in late-type stars

J/A+A/395/11/  table2.dat  (35Kb)

Tosti G., Massaro E., Nesci R., Ciprini S., Nucciarelli G., Maesano M., Montagni F., Raiteri C.M., Villata M., Lanteri L., Ostorero L. Astron. Astrophys. 395, 11 (2002) The optical behaviour of ON 231 (W Comae) during and after the great outburst of spring 1998.

J/A+A/395/97/  table1.dat table1.ps table1.tex  (132Kb)

de Medeiros J.R., Udry S., Burki G., Mayor M. Astron. Astrophys. 395, 97 (2002) A catalog of rotational and radial velocities for evolved stars. II. Ib supergiant stars.

J/A+A/395/129/ table3a.dat table3b.dat table3c.dat table3d.dat table3e.dat table3f.dat table3g.dat table3h.dat table3i.dat table3j.dat table3k.dat table3l.dat table3m.dat table3n.dat table3o.dat (36Kb) Kreykenbohm I., Corburn W., Wilms J., Kretschmar P., Staubert R., Heindl W.A., Rothschild, R.E. Astron. Astrophys. 395, 129 (2002) Confirmation of two cyclotron lines in Vela X-1.

J/A+A/395/347 See I/282/

J/A+A/395/431/  table5.dat table6.dat  (48Kb)

Beuing J., Bender R., Mendes de Oliveira C., Thomas D., Maraston C. Astron. Astrophys. 395, 431 (2002) Line-strength indices and velocity dispersions for 148 early-type galaxies in different environments

J/A+A/395/499/  table1.dat table2.dat table3.dat  (20Kb)

Rauw G., Naze Y., Gosset E., Stevens I.R., Blomme R., Corcoran M.F., Pittard J.M., Runacres M.C. Astron. Astrophys. 395, 499 (2002) An XMM-Newton observation of the Lagoon Nebula and the very young open cluster NGC 6530.

J/A+A/395/587/  tablea1.dat tablea2.dat tablea3.dat tables.ps tables.tex  (79Kb)

Zhang X., Zhang R.X., Fang M.J.

Astron. Astrophys. 395, 587 (2002) Long-term photometry of the near-contact detached binary GR Tauri.

J/A+A/395/753/  table3.dat table4.dat  (37Kb)

Wegner, G., Corsini, E.M., Saglia, R.P., Bender, R., Merkl, D., Thomas, D., Thomas, J., Mehlert, D. Astron. Astrophys. 395, 753 (2002) Spatially resolved spectroscopy of Coma cluster early-type galaxies. II. The minor axis dataset.

J/A+A/395/813/  table3.dat  (21Kb)

Barrado y Navascues D., Bouvier J., Stauffer J.R., Lodieu N., McCaughrean M.J. Astron. Astrophys. 395, 813 (2002) A substellar mass function for Alpha Persei.

J/A+A/395/823/  table4.dat table5.dat table6.dat table7.dat table8.dat  (59Kb)

Paunzen E., Maitzen H.M. Astron. Astrophys. 395, 823 (2002) CCD photometric search for peculiar stars in open clusters. IV. Collinder 272, Pismis 20, Lynga 14, NGC 6396 and NGC 6611

J/A+A/395/885/  table1.dat  (5Kb)

Duemmler R., Iliev I.Kh., Iliev L. Astron. Astrophys. 395, 885 (2002) The radial velocities and physical parameters of HD 553.

J/A+A/396/125/  table2.dat table3.dat  (40Kb)

Frolov V.N., Jilinski E.G., Ananjevskaja J.K., Poljakov E.V., Bronnikova N.M., Gorshanov D.L. Astron. Astrophys. 396, 125 (2002) Proper motions and CCD-photometry of stars in the region of the open cluster NGC 1513.

J/A+A/396/397/  table1.dat  (7Kb)

Boschin W. Astron. Astrophys., 396, 397-409 (2002) A deep cluster survey in Chandra archival data. First results.

J/A+A/396/449/  table4.dat  (29Kb)

Gavazzi G., Boselli A., Pedotti P., Gallazzi A., Carrasco L. Astron. Astrophys. 396, 449 (2002) Hαsurface photometry of galaxies in the Virgo cluster. IV. The current star formation in nearby clusters of galaxies.

J/A+A/396/473/  table1a.dat table1b.dat table1c.dat table1d.dat  (149Kb)

Makarova L.N., Grebel E.K., Karachentsev I.D., Dolphin A.E., Karachentseva V.E., Sharina M.E., Geisler D., Guhathakurta P., Hodge P.W., Sarajedini A., Seitzer P. Astron. Astrophys. 396, 473 (2002) Tidal dwarfs in the M81 group : the second generation?

J/A+A/396/503/  table1.dat table2.dat table3.dat table4.dat table6.dat  (77Kb)

Kong X., Cheng F.Z., Weiss A., Charlot S. Astron. Astrophys. 396, 503 (2002) Spectroscopic study of blue compact galaxies. II. Spectral analysis and correlations.

J/A+A/396/513/  table1.dat table2.dat table3.dat  (203Kb)

Herbst W., Bailer-Jones C.A.L., Mundt R., Meisenheimer K., Wackermann R. Astron. Astrophys. 396, 513 (2002) Stellar rotation and variability in the Orion Nebula Cluster.

J/A+A/396/933/  table1.dat  (17Kb)

Morlet G., Salaman M., Gili R. Astron. Astrophys. 396, 933 (2002) Nice Observatory CCD measurements of visual double stars (4th series).

J/A+A/397/133/  table1.dat table1.ps table3.dat  (1.0 Mb)

Russeil D. Astron. Astrophys. 397, 133 (2003) Star-forming complexes and the spiral structure of our Galaxy.

J/A+A/397/147/  refs.dat stars.dat table1.dat table2.dat  (41Kb)

Pizzolato N., Maggio A., Micela G., Sciortino S., Ventura P. Astron. Astrophys. 397, 147 (2003) The stellar activity-rotation relationship revisited: Dependence of saturated and non-saturated X-ray emission regimes on stellar mass for late-type dwarfs.

J/A+A/397/177/  table1.dat table2.dat  (80Kb)

Bica, E., Dutra C.M., Barbuy B. Astron. Astrophys. 397, 177 (2003) A catalogue of infrared star clusters and stellar groups.

J/A+A/397/213/ subcat1.dat subcat2.dat subcat3.dat subcat4.dat subcat5.dat subcat6.dat subcat7.dat subcat8.dat subcat9.dat subcat10.dat subcat11.dat synthcat.dat (4.5 Mb) Paladini R., Burigana C., Davies R. D., Maino D., Bersanelli M., Cappellini B., Platania P., Smoot G. Astron. Astrophys. 397, 213 (2003) A radio catalog of Galactic HII regions for applications from decimeter to millimeter wavelengths.

J/A+A/397/257/  table3.dat table5.dat  (19Kb)

Liang Y.C., Zhao G., Chen Y.Q., Qiu H.M., Zhang B. Astron. Astrophys. 397, 257 (2003) Chemical compositions of four barium stars.

J/A+A/397/361/  catalog.dat table4.dat table5.dat  (880Kb)

Zhou X., Jiang Z., Ma J., Xue S., Wu H., Chen J., Zhu J., Sun W., Windhorst R.A. Astron. Astrophys. 397, 361 (2003) 15 colour photometry of Landolt SA95 standard star field.

J/A+A/397/421/  a1367.dat coma.dat table2.dat  (896Kb)

Iglesias-Paramo J., Boselli A., Gavazzi G., Cortese L., Vilchez J.M. Astron. Astrophys. 397, 421 (2003) The r’-band luminosity function of Abell 1367: a comparison with Coma.

J/A+A/397/431/  table1.dat  (11Kb)

Mercurio A., Girardi M., Boschin W., Merluzzi P., Busarello G. Astron. Astrophys. 397, 431 (2003) Structure and evolution of galaxy clusters: Internal dynamics of ABCG209 at z 0.21.

J/A+A/397/463/  table2.dat table3.dat  (13Kb)

Ugryumov A.V., Engels D., Pustilnik S.A. Kniazev A.Y., Pramsky A.G., Hagen H.-J. Astron. Astrophys. 397, 463 (2003) The Hamburg/SAO Survey for low metallicity blue compact/HII galaxies (HSS-LM). I. The first list of 46 strong-lined galaxies.

J/A+A/397/473/  table4.dat table5.dat  (103Kb)

Bik A., Lamers H.J.G.L.M., Bastian N., Panagia N., Romaniello M. Astron. Astrophys. 397, 473 (2003) Clusters in the inner spiral arms of M51: The cluster IMF and the formation history.

J/A+A/397/527 See VI/109/

J/A+A/397/575/  lehpms.dat  (1Kb)

Pokorny R.S, Jones H.R.A., Hambly N.C. Astron. Astrophys. 397, 575 (2003) The Liverpool-Edinburgh high proper motion survey

J/A+A/397/711/  table2.dat  (28Kb)

Rau A., Greiner J. Astron. Astrophys. 397, 711 (2003) Comptonization and reflection of X-ray radiation and the X-ray-radio correlation in the χ-states of GRS 1915+105.

J/A+A/397/899/  table1.dat  (26Kb)

Sakamoto T., Chiba M., Beers T.C. Astron. Astrophys. 397, 899 (2003) The mass of the Milky Way: Limits from a newly assembled set of halo objects.

J/A+A/397/921/  table1.dat  (7Kb)

van Genderen A.M. Astron. Astrophys. 397, 921 (2003) The effect of the apsidal motion on the light curve of the close binary HD 93205 (O3V+O8V).

J/A+A/397/997/  table1.dat table2.dat table3.dat table5.dat table7.dat  (263Kb)

Platais I., Pourbaix D., Jorissen A., Makarov V.V., Berdnikov L.N., Samus N.N., Lloyd Evans T., Lebzelter T., Sperauskas J. Astron. Astrophys. 397, 997 (2003) Hipparcos red stars in the HpV_{T 2} and VI_{C} systems.

J/A+A/397/1035/ lines.dat spectra.dat u480490.dat u490500.dat u500510.dat u510520.dat u520530.dat u530540.dat u540550.dat u550560.dat u560570.dat u570580.dat u580590.dat u590600.dat u600610.dat u610620.dat u620630.dat u630640.dat u640650.dat u650660.dat u660670.dat (5.5 Mb) Chentsov E.L., Ermakov S.V., Klochkova V.G., Panchuk V.E., Bjorkman K.S., Miroshnichenko A.S. Astron. Astrophys. 397, 1035 (2003) An atlas of spectra of B6-A2 hypergiants and supergiants from 4800 to 6700Å.

J/A+A/398/L11/  table1.dat  (22Kb)

Thorwirth S., Lichau H. Astron. Astrophys. 398, L11 (2003) The millimeter-wave spectrum and the dipole moment of vinylacetylene.

J/A+A/398/49/  table4.dat  (1.7 Mb)

Heidt J., Appenzeller I., Gabasch A., Jaeger K., Seitz S., Bender R., Boehm A., Snigula J., Fricke K.J., Hopp U., Kuemmel M., Moellenhoff C., Szeifert T., Ziegler B., Drory N., Mehlert D., Moorwood A.F.M., Nicklas H., Noll S., Saglia R.P., Seifert W., Stahl O., Sutorius E., Wagner S.J. Astron. Astrophys. 398, 49 (2003) The FORS Deep Field: Field selection, photometric observations and photometric catalog

J/A+A/398/141/  catalog.dat  (55Kb)

Soubiran C., Bienayme O., Siebert A. Astron. Astrophys. 398, 141 (2003) Vertical distribution of Galactic disk stars. I. Kinematics and metallicity.

J/A+A/398/239/  table1.dat table2.dat  (129Kb)

Ribas I. Astron. Astrophys. 398, 239 (2003) The 0.4-M_{⊙} eclipsing binary CU Cnc: Absolute dimensions, comparison with evolutionary models and possible evidence for a circumstellar dust disk.

J/A+A/398/363/  table2.dat velocity.dat  (11Kb)

Santos N.C., Israelian G., Mayor M., Rebolo R., Udry S. Astron. Astrophys. 398, 363 (2003) Statistical properties of exoplanets. II. Metallicity, orbital parameters, and space velocities

J/A+A/398/479/  notes.dat table1.dat table2.dat  (26Kb)

Karachentsev I.D., Makarov D.I., Sharina M.E., Dolphin A.E., Grebel E.K., Geisler D., Guhathakurta P., Hodge P.W., Karachentseva V.E., Sarajedini A., Seitzer P. Astron. Astrophys. 398, 479 (2003) Local galaxy flows within 5 Mpc.

J/A+A/398/541/  coord.dat daovper.dat daovred.dat ismi.dat ismv.dat  (930Kb)

Kopacki G., Kolaczkowski Z., Pigulski A., Astron. Astrophys. 398, 541 (2003) Variable stars in the globular cluster M13.

J/A+A/398/647/  table3.dat tablea1.dat  (20Kb)

Reiners A., Schmitt J.H.M.M Astron. Astrophys. 398, 647 (2003) Rotation and differential rotation in field F- and G-type stars

J/A+A/398/721/  temp01.dat  (49Kb)

Taylor B.J. Astron. Astrophys. 398, 721 (2003) Statistical cataloging of archival data for class IV-V stars. I. The epoch 2001 temperature catalog.

J/A+A/398/731/  fehd.dat refs.dat  (58Kb)

Taylor B.J. Astron. Astrophys. 398, 731 (2003) Statistical cataloging of archival data for class IV-V stars. II. The epoch 2001 [Fe/H] catalog

J/A+A/398/901/  table1.dat table3.dat table4.dat  (26Kb)

Ciliegi P., Zamorani G., Hasinger G., Lehmann I., Szokoly G., Wilson G. Astron. Astrophys. 398, 901 (2003) A deep VLA survey at 6 cm in the Lockman Hole.

J/A+A/398/1073/  table1.dat  (34Kb)

Carrier F., Barblan F., Burki G., Bartholdi P., Nicolet B. Astron. Astrophys. 398, 1073 (2003) Physical parameters of the Algol system BP Muscae from simultaneous analysis of Geneva 7-colour light curves

J/A+A/398/1121/  table5.dat table6.dat  (41Kb)

Erspamer D., North P. Astron. Astrophys. 398, 1121 (2003) Automated Spectroscopic abundances of A and F-type stars using echelle spectrographs. II. Abundances of 140 A-F stars from ELODIE and CORALIE.

J/A+A/399/105/  position.dat table1.dat table2.dat table3.dat  (38Kb)

Mermilliod J.-C., Latham D.W., Glushkova E.V., Ibrahimov M.A., Batirshinova V.M., Stefanil R.P., James D.J Astron. Astrophys. 399, 105 (2003) Red giants in open clusters: X. NGC 1817

J/A+A/399/141/  table1.dat  (211Kb)

Magakian T.Yu. Astron. Astrophys. 399, 141 (2003) Merged catalogue of reflection nebulae.

J/A+A/399/147/  table5.dat  (71Kb)

Massi F., Giannini T., Lorenzetti D. Astron. Astrophys. 399, 147 (2003) Star formation in the Vela molecular clouds. V. Young stellar objects and star clusters towards the C-cloud.

J/A+A/399/271/  table1.dat table2.dat  (53Kb)

Pinheiro F.J.G., Folha D.F.M., Marconi M., Ripepi V., Palla F., Monteiro M.J.P.F.G., Bernabei S. Astron. Astrophys. 399, 271 (2003) Oscillations in the PMS delta-Scuti star V346 Ori.

J/A+A/399/315/  table1a.dat table1b.dat  (25Kb)

Strassmeier K.G., Rice J.B Astron. Astrophys. 399, 315 (2003) Doppler imaging of stellar surface structure. XIX. The solar-type components of the close binary σ^{2} Coronae Borealis.

J/A+A/399/351 See IX/34/

J/A+A/399/365/  table1.dat  (6Kb)

Blake C., Fox D.W., Park H.S., Williams G.G. Astron. Astrophys. 399, 365 (2003) A search for period changes in δScuti stars with the super-LOTIS sky patrol system.

J/A+A/399/457/  table1.dat table2.dat  (29Kb)

Sanchez-Saavedra M.L., Battaner E., Guijarro A., Lopez-Corredoira M., Castro-Rodriguez N. Astron. Astrophys. 399, 457 (2003) A catalog of warps in spiral and lenticular galaxies in the southern hemisphere.

J/A+A/399/469/  notes.dat tablea1.dat  (13Kb)

Hook I.M., Shaver P.A., Jackson C.A., Wall J.V., Kellermann K.I. Astron. Astrophys. 399, 469 (2003) The Parkes Quarter-Jansky Flat-Spectrum Sample. 2. New optical spectra and redshift measurements.

J/A+A/399/543/  table1.dat table2.dat table4b.dat  (19Kb)

Claria J.J., Piatti A.E., Lapasset E., Mermilliod J.-C. Astron. Astrophys. 399, 543 (2003) Multicolour photometry and Coravel observations of stars in the southern open cluster IC 2488

J/A+A/399/795/  table1.dat table1.ps table1.tex  (70Kb)

Tankosic D., Popovic L.C., Dimitrijevic M.S. Astron. Astrophys. 399, 795 (2003) The electron-impact broadening parameters for Co III.

J/A+A/399/799/  table1.dat table2.dat table3.dat table4.dat tables.ps tables.tex  (988Kb)

Aggarwal K.M., Keenan F.P. Astron. Astrophys. 399, 799 (2003) Effective collision strengths for transitions in Fe XI

J/A+A/399/813/  table10.dat  (18Kb)

Ferrari C., Maurogordato S., Cappi A., Benoist C. Astron. Astrophys. 399, 813 (2003) Multiple merging events in Abell 521

J/A+A/399/919/  tablea1.dat  (12Kb)

Pillitteri I., Micela G., Sciortino S., Favata F. Astron. Astrophys. 399, 919 (2003)

The X-ray Luminosity Distributions of the high-metallicity open cluster Blanco 1.

J/A+A/399/995/  table1.dat table2.dat  (27Kb)

Persi P., Marenzi A.R., Gomez M., Olofsson G. Astron. Astrophys. 399, 995 (2003) Near-infrared and ISOCAM observations of Chamaeleon dark cloud.

J/A+A/399/1083/  table1.dat table2.dat tablea1.dat  (1.3 Mb)

Kerton C.R., Brunt C.M. Astron. Astrophys. 399, 1083 (2003) The association of IRAS sources and ^{12}CO emission in the outer Galaxy.

J/A+A/399/1135/  table3.dat table4.dat  (29Kb)

Deharveng L., Zavagno A., Salas L., Porras A, Caplan J., Cruz-Gonzalez I. Astron. Astrophys. 399, 1135 (2003) Sequential star formation at the periphery of the HII regions Sh 217 and Sh 219.

J/A+A/399/1167/  table1.dat  (18Kb)

Pourbaix D., Platais I., Detournay S., Jorissen A., Knapp G., Makarov V.V. Astron. Astrophys. 399, 1167 (2003) How many Hipparcos Variability-Induced Movers are genuine binaries ?

J/A+A/400/521/  table1.dat table2.dat table3.dat table5.dat  (28Kb)

Gendre B., Barret D., Webb N.A. Astron. Astrophys. 400, 521 (2003) An XMM-Newton observation of the globular cluster Omega Centauri.

J/A+A/400/533/  table3.dat table4.dat  (26Kb)

Dutra, C.M., Bica, E., Soares, J., Barbuy B. Astron. Astrophys. 400, 533 (2003) New infrared star clusters in the southern Milky Way with 2MASS.

J/A+A/400/659/  position.dat table1.dat table3.dat table4.dat tables.ps tables.tex ubvri.dat  (97Kb)

Cutispoto G., Messina S., Rodono M. Astron. Astrophys. 400, 659 (2003) Long-term monitoring of active stars. X. Photometry collected in 1994.

J/A+A/400/791/  table1.dat  (151Kb)

Dimitrijevic M.S., Dacic M., Cvetkovic Z., Sahal-Brechot S. Astron. Astrophys. 400, 791 (2003) Stark broadening of spectral lines of multicharged ions of astrophysical interest. XXIII: Be III spectral lines

J/A+A/400/877/  table1.dat table2.dat  (21Kb)

Pauli E.-M., Napiwotzki R., Altmann M., Heber U., Odenkirchen M., Kerber F. Astron. Astrophys. 400, 877 (2003) 3D kinematics of white dwarfs from the SPY project.

J/A+A/400/891/  table3.dat  (10Kb)

Moraux E., Bouvier J., Stauffer J.R., Cuillandre J.-C. Astron. Astrophys. 400, 891 (2003) Brown dwarfs in the Pleiades cluster: clues to the substellar mass function.

J/A+A/400/903/  table5.dat  (10Kb)

Jiang B.W., Omont A., Ganesh S., Simon G., Schuller F. Astron. Astrophys. 400, 903 (2003) The ISOGAL field FC-01863+00035: Mid-IR interstellar extinction and stellar populations

J/A+A/400/917/  table1.dat table2.dat  (373Kb)

Alcaino G., Alvarado F., Borissova J., Kurtev R. Astron. Astrophys. 400, 917 (2003) The Small Magellanic Cloud star cluster NGC 458. A new UBVI photometric study.

J/A+A/400/939/  obsrun.dat refs.dat table4.dat  (27Kb)

Edelmann H., Heber U., Hagen H.-J., Lemke M., Dreizler S., Napiwotzki R., Engels D. Astron. Astrophys. 400, 939 (2003) Spectral analysis of sdB stars from the Hamburg Quasar Survey.

J/A+A/400/1043/  region.dat table1.dat table2.dat  (103Kb)

Albayrak B., Gulliver A.F., Adelman S.J., Aydin C., Kocer D. Astron. Astrophys. 400, 1043 (2003) A Spectroscopic atlas of Deneb (A2 Iae) λλ3826-5212.

J/A+A/400/1095/  table4.dat  (405Kb)

Veiga C.H., Vieira Martins R., Vienne A., Thuillot W., Arlot J.-E. Astron. Astrophys. 400, 1095 (2003) CCD astrometric observations of Saturnian satellites.

J/A+A/400/1145/  tablea.dat tableb.dat tablec.dat  (317Kb)

Capitaine N., Chapront J., Lambert S., Wallace P.T. Astron. Astrophys. 400, 1145 (2003) Expressions for the Celestial Intermediate Pole and Celestial Ephemeris Origin consistent with the IAU 2000A precession-nutation model.

J/A+A/401/271/  rv.dat vmag.dat  (10Kb)

Carrier F., Burki G. Astron. Astrophys. 401, 271 (2003) Outbursts in the Be star HR 2501.

J/A+A/401/357/  table2.dat table3.dat table4.dat table6.dat tablea.dat  (35Kb)

Adelman S.J. Astron. Astrophys. 401, 357 (2003) uvby Photometry of the Am stars 60 Tau and HR 1528 and the mCP stars HR 8216 and HR 8770

J/A+A/401/471/  table2.dat  (13Kb)

Cortese L., Gavazzi G., Iglesias-Paramo J., Boselli A., Carrasco L. Astron. Astrophys. 401, 471 (2003) Optical spectroscopy and the UV luminosity function of galaxies in the Abell 1367, Coma and Virgo clusters.

J/A+A/401/531/  table8.dat  (254Kb)

Jilinski E.G., Frolov V.N., Ananjevskaja J.K., Straume J., Drake N.A. Astron. Astrophys. 401, 531 (2003) Memberships and CM diagrams of the open cluster NGC 7243.

J/A+A/401/543/  table1.dat table2.dat  (16Kb)

Preibisch T. Astron. Astrophys. 401, 543 (2003) XMM-Newton study of the star forming region NGC 1333.

J/A+A/401/613/  notes.dat refs.dat table1.dat  (14Kb)

Alexander R.D., Casali M.M., Andre P., Persi P., Eiroa C. Astron. Astrophys. 401, 613 (2003) ISOCAM-CVF spectroscopy of the circumstellar environment of young stellar objects.

J/A+A/401/657/ abatlas.dat abphoeni.dat atlas.dat cdatlas.dat cdphoeni.dat efatlas.dat efphoeni.dat original/* phoenix.dat sun.dat uatlas.dat uphoenix.dat vega.dat (43.3 Mb) Claret A. Astron. Astrophys. 401, 657 (2003) A new non-linear limb-darkening law for LTE stellar atmosphere models. II. Geneva and Walraven systems: Calculations for -5.0 =log[M/H] =+1, 2000K =T_{eff} =50000K at several surface gravities.

J/A+A/401/873/  tablea.dat tableb.dat tablec.dat tabled.dat tablee.dat  (135Kb)

Wilke K., Stickel M., Haas M., Herbstmeier U., Klaas U., Lemke D. Astron. Astrophys. 401, 873 (2003) The Small Magellanic Cloud in the far infrared. I. ISO’s 170μm map and revisit of the IRAS 12-100μm data.

J/A+A/401/927/  table8.dat table9.dat  (32Kb)

Beckmann V., Engels D., Bade N., Wucknitz O. Astron. Astrophys. 401, 927 (2003) The HRX-BL Lac sample - Evolution of BL Lac objects.

J/A+A/401/959/ table2a.dat table2b.dat table2c.dat table3a.dat table3b.dat table3c.dat table5a.dat table5b.dat (28Kb) Phan-Bao N., Crifo F., Delfosse X., Forveille T., Guibert J., Borsenberger J., Epchtein N., Fouque P., Simon G., Vetois J. Astron. Astrophys., 401, 959-974 (2003) New neighbours. V. 35 DENIS late-M dwarfs between 10 and 30 parsecs.

J/A+A/401/997/  table1.dat  (12Kb)

Pace G., Pasquini L., Ortolani S. Astron. Astrophys. 401, 997 (2003) The Wilson-Bappu effect: A tool to determine stellar distances.

J/A+A/402/1/  table1a.dat table1b.dat table1c.dat table1d.dat  (38Kb)

Chandra S.

Astron. Astrophys. 402, 1 (2003) Transfer of radiation through cyclopropenylidene and ethylene oxide.

J/A+A/402/37/  table10.dat table11.dat table12.dat  (198Kb)

Boselli A., Gavazzi G., Sanvito G. Astron. Astrophys. 402, 37 (2003) UV to radio centimetric spectral energy distributions of optically-selected late-type galaxies in the Virgo cluster.

J/A+A/402/113/  lc/* table2.dat table4.dat table5.dat  (202Kb)

Joshi Y.C., Pandey A.K., Narashimha D., Sagar R., Giraud-Heraud Y. Astron. Astrophys. 402, 113 (2003) Identification of 13 Cepheids and 333 other Variables in M31.

J/A+A/402/151/  table3.dat table4.dat table5.dat table6.dat table7.dat  (168Kb)

Raiteri C.M., Villata M., Tosti G., Nesci R., Massaro E., Aller M.F., Aller H.D., Terasranta H., Kurtanidze O.M., Nikolashvili M.G., Ibrahimov M.A., Papadakis I.E., Krichbaum T.P., Kraus A., Witzel A., Ungerechts H., Lisenfeld U., Bach U., Cimo G., Ciprini S., Fuhrmann L., Kimeridze G.N., Lanteri L., Maesano M., Montagni F., Nucciarelli G., Ostorero L. Astron. Astrophys. 402, 151 (2003) Optical and radio behaviour of the BL Lacertae object 0716+714.

J/A+A/402/171/  table3.dat  (7Kb)

Tschager W., Schilizzi R.T., Roettgering H.J.A., Snellen I.A.G., Miley G.K., Perley R.A. Astron. Astrophys. 402, 171 (2003) The determination of the spectral turnover in faint compact steep-spectrum radio sources using the VLA at 74 MHz.

J/A+A/402/343/  stars.dat table2.dat table3.dat table5.dat table7.dat  (382Kb)

Takeda Y. Astron. Astrophys. 402, 343 (2003) Oxygen line formation in late-F through early-K dwarfs/giants: Infrared O I triplet and [O I] lines.

J/A+A/402/425/  models.dat  (14Kb)

Mouhcine M., Lancon A. Astron. Astrophys. 402, 425 (2003) Photometric properties of Single Stellar Populations: From Z/Z_{⊙}=1/50 to Z/Z_{⊙}=2.5.

J/A+A/402/433 See III/232/

J/A+A/402/509/ summary.dat table25.dat table26.dat table27.dat table28.dat table29.dat table30.dat table31.dat table32.dat table33.dat table34.dat table35.dat table36.dat table37.dat table38.dat table39.dat table40.dat (132Kb) Clausen J.V., Storm J., Larsen S.S., Gimenez A. Astron. Astrophys. 402, 509 (2003) Eclipsing binaries in the Magellanic Clouds. uvby CCD light curves and photometric analyses for HV982 (LMC), HV12578 (LMC), HV1433 (SMC), and HV11284 (SMC).

J/A+A/402/549/  table1.dat  (149Kb)

Baume G., Vazquez R.A., Carraro G., Feinstein A. Astron. Astrophys. 402, 549 (2003) Photometric study of the young open cluster NGC 3293.

J/A+A/402/571/  notes.dat sources.dat tables.tex x-obs.dat  (67Kb)

Huensch M., Weidner C., Schmitt J.H.M.M. Astron. Astrophys., 402, 571-586 (2003) An X-ray study of the open clusters NGC 2451 A and B.

J/A+A/402/647/  tablea1.dat  (19Kb)

Cusumano G., Massaro E. Astron. Astrophys. 402, 647 (2003) Timing noise, glitches and the braking index of PSR B0540-69.

J/A+A/402/745/  table1.dat  (18Kb)

Duemmler R., Doucet C., Formanek F., Ilyin I., Tuominen I. Astron. Astrophys. 402, 745 (2003) The radial velocities and physical parameters of ER Vul.

J/A+A/402/963/  refs.dat table2.dat table2.ps table2.tex  (107Kb)

Maheswar G., Manoj P., Bhatt H.C. Astron. Astrophys. 402, 963 (2003)

Optical spectroscopic and 2MASS measurements of Stephenson Hαstars.

J/A+A/402/1033/  table2.dat table3.dat  (70Kb)

Aarum Ulvas V., Henry G.W. Astron. Astrophys. 402, 1033 (2003) BV photometry of UX Ari in the period 1987-2002.

J/A+A/402/1157/  table1.dat table2.dat table3.dat tables.ps tables.tex  (73Kb)

Nakos T., Ofek E.O., Boumis P., Cuypers J., Sinachopoulos D., van Dessel E., Gal-Yam A., Papamastorakis J. Astron. Astrophys. 402, 1157 (2003) A catalog of secondary photometric standard stars around gravitational lenses.

J/A+A/403/93/  ngc147.dat ngc185.dat  (2.3 Mb)

Nowotny W., Kerschbaum F., Olofsson H., Schwarz H.E. Astron. Astrophys. 403, 93 (2003) A census of AGB stars in Local Group galaxies. II. NGC 185 and NGC 147.

J/A+A/403/141/  table1.dat table2a.dat table2b.dat  (536Kb)

Ojha D.K., Omont A., Schuller F., Simon G., Ganesh S., Schultheis M. Astron. Astrophys. 403, 141 (2003) Stellar sources in the ISOGAL intermediate bulge fields.

J/A+A/403/247/  table7.dat  (120Kb)

Fuhrmeister, B., Schmitt, J.H.M.M. Astron. Astrophys. 403, 247 (2003) A systematic study of X-ray variability in the ROSAT All-Sky Survey.

J/A+A/403/493/  table2.dat  (434Kb)

Moy E., Barmby P., Rigopoulou D., Huang J.-S., Willner S.P., Fazio G.G. Astron. Astrophys. 403, 493 (2003) H-band observations of the Chandra Deep Field South.

J/A+A/403/555/  table1.dat table2.dat  (62Kb)

Artyukh V.S., Kopylov A.I., Kopylova F.G. Astron. Astrophys. 403, 555 (2003) Investigation of radio sources from the Interplanetary Scintillation Pushchino Survey. I. Cross-identification with the 7C and FIRST radio catalogues.

J/A+A/403/637/  table2a.dat table2b.dat table2c.dat  (52Kb)

Csizmadia Sz., Biro I B , Borkovits T. Astron. Astrophys. 403, 637 (2003) Photometric Study of the Over-Contact Binary Star GSC 3822-1056.

J/A+A/403/659/  position.dat refs.dat table4.dat  (33Kb)

Acker A., Neiner C. Astron. Astrophys. 403, 659 (2003) Quantitative classification of WR nuclei of planetary nebulae.

J/A+A/403/675/  bmag.dat umag.dat vmag.dat  (136Kb)

Ozdemir S., Mayer P., Drechsel H., Demircan O., Ak H. Astron. Astrophys. 403, 675 (2003) Refinement of third body parameters and new photometric results for the early-type multiple system IU Aurigae.

J/A+A/403/857/  catalog.dat  (116Kb)

Bondi M., Ciliegi P., Zamorani G., Gregorini L., Vettolani G., Parma P., de Ruiter H., Le Fevre O., Arnaboldi M., Guzzo L., Maccagni D., Scaramella R., Adami C., Bardelli S., Bolzonella M., Bottini D., Cappi A., Foucaud S., Franzetti P., Garilli B., Gwyn S., Ilbert O., Iovino A., Le Brun V., Marano B., Marinoni C., McCracken H.J., Meneux B., Pollo A., Pozzetti L., Radovich M., Ripepi V., Rizzo D., Scodeggio M., Tresse L., Zanichelli A., Zucca E. Astron. Astrophys. 403, 857 (2003) The VLA-VIRMOS Deep Field. I. Radio observations probing the μJy source population.

J/A+A/403/937/  stars.dat table3.dat table4.dat  (113Kb)

Paunzen E., Maitzen H.M., Rakos K.D., Schombert J. Astron. Astrophys. 403, 937 (2003) Stroemgren uvby photometry of the open clusters NGC 6192 and NGC 6451.

J/A+A/403/943/  table3.dat table3s.dat table4.dat table4s.dat  (139Kb)

Le Bertre T., Tanaka M., Yamamura Y., Murakami H.

Astron. Astrophys. 403, 943 (2003) Galactic mass-losing AGB stars probed with the IRTS. II.

J/A+A/403/955 See II/243/

J/A+A/403/993/  table1.dat  (23Kb)

Knapp G.R., Pourbaix D., Platais I., Jorissen A. Astron. Astrophys. 403, 993 (2003) Reprocessing the Hipparcos data of evolved stars. III. Revised Hipparcos period-luminosity relationship for galactic long-period variable stars.

J/A+A/403/1095/  table1.dat  (30Kb)

Minier V., Ellingsen S.P., Norris R.P., Booth R.S. Astron. Astrophys. 403, 1095 (2003) The protostellar mass limit for 6.7 GHz methanol masers. I. A low-mass YSO survey.

J/A+A/403/1105/  table5.dat  (30Kb)

Francois P., Depagne E., Hill V., Spite M., Spite F., Plez B., Beers T.C., Barbuy B., Cayrel R., Andersen J., Bonifacio P., Molaro P., Nordstroem F., Primas B. Astron. Astrophys. 403, 1105 (2003) First stars. III. A detailed elemental abundance study of four extremely metal-poor giant stars

J/A+A/404/L29/  table2.dat table3.dat table4.dat  (10Kb)

Paardekooper S.J., Van der Hucht K.A, van Genderen A., Brogt E., Gieles M., Meijerink R. Astron. Astrophys. 404, L29 (2003) New type of brightness variations of the colliding wind WO4+O5((f)) binary WR30a

J/A+A/404/187/  stars.dat table1.dat table2.dat table4.dat table5.dat table67.dat  (269Kb)

Gratton R.G., Carretta E., Claudi R., Lucatello S., Barbieri M. Astron. Astrophys. 404, 187 (2003) Abundances for metal-poor stars with accurate parallaxes. I. Basic data.

J/A+A/404/223/  table1.dat table2.dat  (29Kb)

Bica E., Dutra C.M., Soares J., Barbuy B. Astron. Astrophys. 404, 223 (2003) New infrared star clusters in the Northern and Equatorial Milky Way with 2MASS.

J/A+A/404/301 See V/113D/

J/A+A/404/423/  table1.dat  (31Kb)

Tammann G.A., Sandage A., Reindl, B. Astron. Astrophys. 404, 423 (2003) New Period-Luminosity and Period-Color Relations of Classical Cepheids: I. Cepheids in the Galaxy,

J/A+A/404/545/  spectrum.dat  (22Kb)

Zhang Y., Liu X.-W. Astron. Astrophys. 404, 545 (2003) Optical spectrum of the planetary nebula M 2-24.

J/A+A/404/621/ list.dat table3.tex table3a.dat table3b.dat table3c.dat table3d.dat table3e.dat table3f.dat table3g.dat table3h.dat table3i.dat table3j.dat table3k.dat table3l.dat table3m.dat table3n.dat table3o.dat table3p.dat table3q.dat table3r.dat table3s.dat table3t.dat table3u.dat (279Kb) Jin H., Kim S.-L., Kwon S.-G., Youn J.H., Lee C.-U., Lee D.-J., Kim K.-S. Astron. Astrophys. 404, 621 (2003) Multiband photometric re-classification of ROTSE-I delta Scuti type stars.

J/A+A/404/661/ stars.dat tablea1.dat tablea2.dat tablea3.dat tablea4.dat tablea5.dat tablea6.dat tablea7.dat tableb1.dat tableb2.dat tableb3.dat (44Kb) Barbuy B., Perrin M.-N., Katz D., Coelho P., Cayrel R., Spite M., van’t Veer Menneret C. Astron. Astrophys. 404, 661 (2003) A grid of synthetic spectra and indices Fe5270, Fe5335, Mgb and Mg_{2} as a function of stellar parameters and [alpha/Fe].

J/A+A/404/689/  stars.dat table3.dat table4.dat table5.dat table6.dat table7.dat  (35Kb)

Niemczura E. Astron. Astrophys. 404, 689 (2003) Metallicities of the SPB stars from the IUE ultraviolet spectra.

J/A+A/404/775/  table4.dat  (7Kb)

Zucker, S., Mazeh, T., Santos, N.C., Udry, S., Mayor, M. Astron. Astrophys. 404, 775 (2003) Multi-order TODCOR: application to observations taken with the CORALIE echelle spectrograph. I. The

system HD 41004.

J/A+A/404/913/  refs.dat table1.dat table2.dat table3.dat table4.dat  (203Kb)

Sartori M.J., Lepine J.R.D., Dias W.S. Astron. Astrophys. 404, 913 (2003) Formation scenarios for the young stellar associations between galactic longitudes l=280^{∘}-360^{∘}.

J/A+A/404/927/  table4.dat  (163Kb)

Prisinzano L., Micela G., Sciortino S., Favata F. Astron. Astrophys. 404, 927 (2003) Luminosity and Mass Function of the Galactic open cluster NGC 2422.

J/A+A/405/1 See VII/234/

J/A+A/405/5/  catalog.dat refs.dat  (331Kb)

Bettoni D., Galletta G., Garcia-Burillo S. Astron. Astrophys. 405, 5 (2003) A new catalogue of ISM content of normal galaxies

J/A+A/405/99/  fig2a.ps fig2b.ps fig2c.ps fig2d.ps table1.dat table2.dat table3.dat  (4.1 Mb)

Monnier-Ragaigne D., van Driel W., Schneider S.E., Jarrett T.H., Balkowski C. Astron. Astrophys. 405, 99 (2003) A search for Low Surface Brightness galaxies in the near-infrared. I. Selection of the sample.

J/A+A/405/111/  catalog1.dat catalog2.dat table3.dat  (131Kb)

Gouliermis D., Kontizas M., Kontizas E., Korakitis R. Astron. Astrophys. 405, 111 (2003) OB stellar associations in the Large Magellanic Cloud: Survey of young stellar systems.

J/A+A/405/231/ abper.dat acuma.dat cpand.dat rscep.dat rvper.dat rwcrb.dat rwper.dat rxhya.dat stars.dat stper.dat twdra.dat vzleo.dat xtri.dat xzuma.dat zdra.dat (86Kb) Kim S.-L., Lee J.W., Kwon S.-G., Youn J.-H., Mkrtichian D.E., Kim C. Astron. Astrophys. 405, 231 (2003) Search for A-F spectral type pulsating components in Algol-type eclipsing binary systems.

J/A+A/405/531/  spectra/* table1.dat  (3.2 Mb)

Schultheis M., Lancon A., Omont A., Schuller F., Ojha D.K. Astron. Astrophys. 405, 531 (2003) Near-IR spectra of ISOGAL sources in the inner Galactic Bulge.

J/A+A/405/627/  table1.dat table2.dat  (57Kb)

Tylenda R., Siodmiak N., Gorny S.K., Corradi R.L.M., Schwarz H.E. Astron. Astrophys. 405, 627 (2003) Angular dimension of planetary nebulae

J/A+A/405/999/  table1.dat table2.dat  (12Kb)

Kahanpaeae J., Mattila K., Lehtinen K., Leinert C., Lemke D. Astron. Astrophys. 405, 999 (2003) Unidentified infrared bands in the interstellar medium across the Galaxy.

J/A+A/405/1087/  table1.dat table2.dat  (35Kb)

Yakut K., Tarasov A.E., Ibanoglu C., Harmanec P., Kalomeni B., Holmgren D.E., Bozic H., Eenens P. Astron. Astrophys. 405, 1087 (2003) Basic physical properties of the close binary V497 Cephei in the open cluster NGC 7160.

J/A+A/405/1095/  cgm.dat mlt050.dat mlt125.dat tables.tex  (567Kb)

Barban C., Goupil M.J., Van’t Veer-Menneret C., Garrido R., Kupka F., Heiter U. Astron. Astrophys. 405, 1095 (2003) New grids of ATLAS9 atmospheres. II. Limb-darkening coefficients for the Stroemgren photometric system for A-F stars.

J/A+A/406/L23/  rv.dat  (100Kb)

Carrier F., Bourban G. Astron. Astrophys. 406, L23 (2003) Solar-like oscillations in the K1 dwarf star alpha Cen B

J/A+A/406/L47/  freq17oh.dat  (17Kb)

Polehampton E.T., Brown J.M., Swinyard B.M., Baluteau J.-P. Astron. Astrophys. 406, L47 (2003) Far-infrared detection of ^{17}OH towards Sagittarius B2.

J/A+A/406/L9/  table4.dat  (14Kb)

Aharonian F., Akhperjanian A., Beilicke M., Bernloehr K., Boerst H.-G., Bojahr H., Bolz O., Coarasa T., Contreras J.L., Cortina J., Denninghoff S., , Fonseca M.V., Girma M., Goetting N., Heinzelmann G., Hermann G., Heusler A., Hofmann W., Horns D., Jung I., Kankanyan R., Kestel M., Kohnle A., Konopelko A., Kornmeyer H., Kranich D., Lampeitl H., Lopez M., Lorenz E., Lucarelli F., Mang O., Meyer H., Mirzoyan R., Moralejo A., Ona-Wilhelmi E., Panter M., Plyasheshnikov A., Puehlhofer G., delosReyes R., Rhode W., Ripken J., Robrade J., Rowell G., Sahakian V., Samorski M., Schilling M., Siems M., Sobzynska D., Stamm W., Tluczykont M., Vitale V., Voelk H.J., Wiedner C.A., Wittek W. Astron. Astrophys. 406, L9 (2003) Detection of TeV Gamma-Rays from the BL Lac 1ES1959+650 in its low states and during a major outburst in 2002.

J/A+A/406/51/  table1.dat table4.dat  (88Kb)

Cioni M.-R.L., Blommaert J.A.D.L., Groenewegen M.A.T., Habing H.J., Hron J., Kerschbaum F., Loup C., Omont A., van Loon J.T., Whitelock P.A., Zijlstra A.A. Astron. Astrophys. 406, 51 (2003) Long period variables detected by ISO in the Small Magellanic Cloud.

J/A+A/406/75/ table1.dat table2.dat table3.dat table4.dat table5.dat table6.dat table7.dat table8.dat (1.5 Mb) Rejkuba M., Minniti D., Silva D.R. Astron. Astrophys. 406, 75 (2003) Long period variables in NGC 5128. I. Catalogue.

J/A+A/406/193/  bmag.dat rmag.dat vmag.dat  (60Kb)

Zboril M. Astron. Astrophys. 406, 193 (2003) SV Cam spot activity in February 2001 - March 2002.

J/A+A/406/373/  table1.dat table2.dat  (15Kb)

Santos N.C., Udry S., Mayor M., Naef D., Pepe F., Queloz D., Burki G., Cramer N., Nicolet B. Astron. Astrophys. 406, 373 (2003) The CORALIE survey for southern extra-solar planets. XI. The return of the giant planet orbiting HD 192263.

J/A+A/406/535/  optical.dat x-ray.dat  (1.9 Mb)

Zickgraf F.-J., Engels D., Hagen H.-J., Reimers D., Voges W. Astron. Astrophys. 406, 535 (2003) The Hamburg/RASS Catalogue of optical identifications. Northern high-galactic latitude ROSAT Bright Source Catalogue X-ray sources.

J/A+A/406/579/  table3.dat table4.dat  (46Kb)

Klein U., Mack K.-H., Gregorini L., Vigotti M. Astron. Astrophys. 406, 579 (2003) Multi-frequency study of the B3-VLA sample. III. Polarisation properties.

J/A+A/406/593/  table1.dat table3.dat  (151Kb)

Baker A.J., Davies R.I., Lehnert M.D., Thatte N.A., Vacca W.D., Hainaut O.R., Jarvis, M.J., Miley G.K., Roettgering H.J.A. Astron. Astrophys. 406, 593 (2003) Galaxies in Southern Bright Star Fields I. Near-infrared imaging

J/A+A/406/751/  tablea1.dat tablea2.dat tablea3.dat tablea4.dat  (167Kb)

Fienga A., Bange J.F., Bec-Borsenberger A., Thuillot W. Astron. Astrophys. 406, 751 (2003) Close encounters of asteroids before and during the ESA GAIA mission.

J/A+A/406/829/  wsrtgals.dat  (18Kb)

Braun R., Thilker D., Walterbos R.A.M. Astron. Astrophys. 406, 829 (2003) The WSRT wide-field HI survey: I The background galaxy sample

J/A+A/406/867/  table1.dat table2.dat  (42Kb)

Boselli A., Sauvage M., Lequeux J., Donati A., Gavazzi G. Astron. Astrophys. 406, 867 (2003) Mid-IR emission of galaxies in the Virgo cluster. III. The data.

J/A+A/406/975/  abund.dat table5.dat table6.dat tables.ps tables.tex  (127Kb)

Kocer D., Adelman S.J., Caliskan H., Gulliver A.F., Tektunali. G.H. Astron. Astrophys. 406, 975 (2003) Elemental abundance analyses with DAO spectrograms. XXVII. The superficially normal stars theta And (A2 IV), epsilon Del (B6 III), epsilon Aqr (A1.5 V), and iota And (B9 V).

J/A+A/406/987/  abund.dat table5.dat table6.dat table7.dat tables.ps tables.tex  (284Kb)

Pintado O.I., Adelman S.J. Astron. Astrophys. 406, 987 (2003) Elemental abundance analyses with the EBASIM spectrograph of the 2.1-m CASLEO Observatory Telescope. I. The late B and early A stars xi Octantis, alpha Sextantis, and 68 Taurus.

J/A+A/406/995/  refs.dat spectra/* table1.dat  (2.6 Mb)

Marrese P.M., Boschi F., Munari U. Astron. Astrophys. 406, 995 (2003) High resolution spectroscopy over λλ8500-8750Åfor GAIA. IV. Extending the cool MK stars sample.

J/A+A/406/1135/  tablea.dat tableb.dat tablec.dat tablef.dat  (340Kb)

Capitaine N., Wallace P.T., McCarthy D.D. Astron. Astrophys. 406, 1135 (2003) Expressions to implement the IAU 2000 definition of UT1.

J/A+A/407/31/  table1.dat  (12Kb)

Proust D., Capelato H.V., Hickel G., Sodre Jr. L., Lima Neto G.B., Cuevas H. Astron. Astrophys. 407, 31 (2003) The cluster of galaxies Abell 376.

J/A+A/407/423/  position.dat tablea1.dat tablea2.dat tablea3.dat  (20Kb)

Mehlert D., Thomas D., Saglia R.P., Bender R., Wegner G. Astron. Astrophys. 407, 423 (2003) Spatially resolved spectroscopy of Coma cluster early-type galaxies: III. The stellar population gradients.

J/A+A/407/527/  table3.dat table5.dat  (269Kb)

Baume G., Villanova S., Carraro G. Astron. Astrophys. 407, 527 (2003) A study of the two northern open clusters NGC 1582 and NGC 1663.

J/A+A/407/631/  refs.dat tablea1.dat  (85Kb)

Bychkov V.D., Bychkova L.V., Madej J. Astron. Astrophys. 407, 631 (2003) Catalogue of averaged stellar effective magnetic fields. I. Chemically peculiar A and B type stars.

J/A+A/407/769/  table1.dat table34.dat tables.tex  (301Kb)

Aggarwal K.M., Keenan F.P. Astron. Astrophys. 407, 769 (2003) Excitation rates for transitions in Ca XV.

J/A+A/407/791/  table4.dat table5.dat table6.dat table7.dat table8.dat table9.dat  (27Kb)

Metcalfe L., Kneib J.-P., McBreen B., Altieri B., Biviano A., Delaney M., Elbaz D., Kessler M.F., Leech K., Okumura K., Ott S., Perez-Martinez R., Sanchez-Fernandez C., Schulz B. Astron. Astrophys. 407, 791 (2003) An ISOCAM survey through gravitationally lensing galaxy clusters. I. Source lists and source counts for A370, A2218 and A2390.

J/A+A/407/899/  table1.dat  (28Kb)

Fan J.H., Zhang J.S. Astron. Astrophys. 407, 899 (2003) The core dominance parameter of extragalactic radio sources.

J/A+A/407/919/  table1.dat table4.dat  (180Kb)

Alcaino G., Alvarado F., Kurtev R. Astron. Astrophys. 407, 919 (2003) The SMC star cluster Lindsay 1. First UI photometric study.

J/A+A/407/1039/  table2.dat table2.fit  (325Kb)

Pottschmidt K., Wilms J., Nowak M.A., Pooley G.G., Gleissner T., Heindl W.A., Smith D.M., Remillard R., Staubert R. Astron. Astrophys. 407, 1039 (2003) Long term variability of Cygnus X-1 I. X-ray spectral-temporal correlations in the hard state.

J/A+A/407/1059/  table2.dat table5.dat table7.dat tables.ps tables.tex  (58Kb)

Rodriguez E., Arellano Ferro A., Costa V., Lopez-Gonzalez M.J., Sareyan J.P. Astron. Astrophys. 407, 1059 (2003) Delta Sct-type nature of the variable V2109 Cyg.

J/A+A/407/1139/  table3.dat table4.dat table5.dat tables.ps tables.tex  (61Kb)

Rousselot P., Petit J.-M., Poulet F., Lacerda P., Ortiz J. Astron. Astrophys. 407, 1139 (2003)

Photometry of the Kuiper-Belt object 1999 TD_{10} at different phase angles.

J/A+A/407/1157/  table4.dat table5.dat table6.dat table7.dat table8.dat table9.dat  (118Kb)

Hanuschik R.W. Astron. Astrophys. 407, 1157 (2003) A flux-calibrated, high-resolution atlas of optical sky emission from UVES.

J/A+A/408/L29/  table2.dat  (9Kb)

Arellano Ferro A., Giridhar S. Astron. Astrophys. 408, L29 (2003) The new T Tauri double-lined spectroscopic binary HD 34700.

J/A+A/408/67/ fig2a.ps fig2b.ps fig2c.ps fig2d.ps fig2e.ps fig3.ps table2.dat table3a.dat table3b.dat table4a.dat table4b.dat table5a.dat table5b.dat (2.3 Mb) Monnier-Ragaigne D., van Driel W., O’Neil K., Schneider S.E., Balkowski C., Jarrett T.H. Astron. Astrophys. 408, 67 (2003) A search for Low Surface Brightness galaxies in the near-infrared. II. Arecibo HI line observations.

J/A+A/408/305/  table1.dat table2.dat  (14Kb)

Miroshnichenko A.S., Bjorkman K.S., Morrison N.D., Wisniewski J.P., Manset N., Levato H., Grosso M., Pollmann E., Buil C., Knauth D.C. Astron. Astrophys. 408, 305 (2003) Spectroscopy of the growing circumstellar disk in the δScorpii Be binary.

J/A+A/408/387/  tables.dat tables.tex  (52Kb)

Bretagnon P., Brumberg V.A. Astron. Astrophys. 408, 387 (2003) On transformation between International Celestial and Terrestrial Reference systems.

J/A+A/408/465/ fig1a.ps fig1b.ps fig1c.ps fig1d.ps fig1e.ps fig1f.ps fig1g.ps fig1h.ps fig2a.ps table3a.dat table3b.dat table4a.dat table4b.dat table5a.dat table5b.dat (3.7 Mb) Monnier Ragaigne D., van Driel W., Schneider S.E., Balkowski C., Jarrett T.H. Astron. Astrophys. 408, 465 (2003) A search for Low Surface Brightness galaxies in the near-infrared. III. Nancay HI line observations.

J/A+A/408/611/  refs.dat table4.dat  (102Kb)

Janik J., Harmanec P., Lehmann H., Yang S., Bozic H., Ak H., Hadrava P., Eenens P., Ruzdjak D., Sudar D., Hubeny I., Linnell A.P. Astron. Astrophys. 408, 611 (2003) Search for forced oscillations in binaries. IV. The eclipsing binary V436 Per revisited.

J/A+A/408/789/ table2.txt table2a.dat table2b.dat table4.dat table4.txt table5.dat table5.txt table7.txt table7a.dat table7b.dat table7c.dat (4.4 Mb) Nahar S.N., Eissner W., Chen G.X., Pradhan A.K. Astron. Astrophys. 408, 789 (2003) Atomic data from the Iron Project. LIII. Relativistic allowed and forbidden transition probabilities for FeXVII.

J/A+A/408/873/  table3.dat table4.dat  (54Kb)

Corsini E.M., Pizzella A., Coccato L., Bertola F. Astron. Astrophys. 408, 873 (2003) Minor-axis velocity gradients in spirals and the case of inner polar disks

J/A+A/408/905/  table1.dat table2.dat tableb.dat  (32Kb)

Cappi A., Benoist C., da Costa L.N., Maurogordato S. Astron. Astrophys. 408, 905 (2003) Nature and environment of Very Luminous Galaxies.

J/A+A/408/1029/  table1.dat table1.tsv  (244Kb)

Kerber F., Mignani R.P., Guglielmetti F., Wicenec A. Astron. Astrophys. 408, 1029 (2003) Galactic Planetary Nebulae and their central stars. I. An accurate and homogeneous set of coordinates.

J/A+A/409/79/  table1.dat  (19Kb)

Fiore F., Brusa M., Cocchia F., Baldi A., Carangelo N., Ciliegi P., Comastri A., La Franca F., Maiolino R., Matt G., Molendi S., Mignoli M., Perola G.C., Severgnini P., Vignali C. Astron. Astrophys. 409, 79 (2003) The HELLAS2XMM survey. IV. Optical identifications and the evolution of the accretion luminosity in the Universe.

J/A+A/409/115/  table1.dat  (15Kb)

Nagar N.M., Wilson A.S., Falcke H., Veilleux S., Maiolino R.

Astron. Astrophys. 409, 115 (2003) The AGN content of ultraluminous IR galaxies: High resolution VLA imaging of the IRAS 1 Jy ULIRG sample.

J/A+A/409/147/  table1.dat  (47Kb)

Preibisch T., Stanke T., Zinnecker H. Astron. Astrophys. 409, 147 (2003) Constraints on the IMF and the brown dwarf population of the young cluster IC 348.

J/A+A/409/169/  ubvri.dat  (38Kb)

Bouvier J., Grankin K.N., Alencar S.H.P., Dougados C., Fernandez M., Basri G., Batalha C., Guenther E., Ibrahimov M.A., Magakian T.Y., Melnikov S.Y., Petrov P.P., Rud M.V., Zapatero Osorio M.R. Astron. Astrophys., 409, 169-192 (2003) Eclipses by circumstellar material in the T Tauri star AA Tau. II. Evidence for non-stationary magnetospheric accretion.

J/A+A/409/245/  table1.dat table2.dat table3.dat  (14Kb)

Tokovinin A., Balega Y.Y., Pluzhnik E.A., Shatsky N.I., Gorynya N.A., Weigelt G. Astron. Astrophys. 409, 245 (2003) Fundamental parameters and origin of the very eccentric binary 41 Dra

J/A+A/409/251/  stars.dat table1.dat  (25Kb)

Mallik S.V., Parthasarathy M., Pati A. Astron. Astrophys. 409, 251 (2003) Lithium and rotation in F and G dwarfs and subgiants

J/A+A/409/275/  table7.dat  (7Kb)

Neiner C., Hubert A.M., Fremat Y., Floquet M., Jankov S., Preuss O., Henrichs H.F., Zorec J. Astron. Astrophys. 409, 275 (2003) Rotation and magnetic field in the Be star omega Orionis

J/A+A/409/361/  catalog.dat table1.dat  (90.2 Mb)

Camargo J.I.B., Ducourant C., Teixeira R., Le Campion J.-F., Rapaport M., Benevides-Soares P. Astron. Astrophys. 409, 361 (2003) Extension of the ICRF for selected areas down to the 16th magnitude - II.

J/A+A/409/439/  table1.dat table2.dat tablea.dat  (33Kb)

Olsen L.F., Hansen L., Jorgensen H.E., Benoist C., da Costa L., Scodeggio M. Astron. Astrophys. 409, 439 (2003) Multi-object spectroscopy of low redshift EIS clusters. II.

J/A+A/409/541/  cross.dat table1.dat  (52Kb)

Claria J.J., Lapasset, E., Piatti, A.E., Ahumada, A.V. Astron. Astrophys. 409, 541 (2003) IC 2395 and BH 47: only one open cluster in the Vela constellation

J/A+A/409/857/  table2.dat  (10Kb)

Andruchow I., Cellone S.A., Romero G.E., Dominici T.P., Abraham Z. Astron. Astrophys. 409, 857 (2003) Microvariability in the optical polarization of 3C 279.

J/A+A/409/933/  table1.dat table2.dat  (14Kb)

Feinstein C., Baume G., Vergne M.M., Vazquez R.A. Astron. Astrophys. 409, 933 (2003) Optical polarization observations in the region of Stock 16.

J/A+A/409/959/  table1.dat  (46Kb)

Degirmenci O.L., Gulmen O., Sezer C., Ibanoglu C., Cakirli O. Astron. Astrophys. 409, 959 (2003) Up-to-date UBV light and O-C curves analyses of the eclipsing binary V477 Cygni

J/A+A/409/1017/  table2.dat table3.dat table4.dat table5.dat table6.dat table7.dat table8.dat  (15Kb)

Messina S., Guinan E.F. Astron. Astrophys. 409, 1017 (2003) Magnetic activity of six young solar analogues. II. Surface Differential Rotation from long-term photometry

J/A+A/409/1037/  table3a.dat table3b.dat  (102Kb)

Alencar S.H.P., Melo C.H.F., Dullemond C.P., Andersen J., Batalha C., Vaz L.P.R., Mathieu R.D. Astron. Astrophys. 409, 1037 (2003) The pre-main sequence spectroscopic binary AK Scorpii revisited.

J/A+A/410/33/  tablea2.dat  (9Kb)

Menard B., Peroux C. Astron. Astrophys. 410, 33 (2003) Investigating lensing by absorbers in the 2dF-quasar survey.

J/A+A/410/53/  tablea1.dat  (18Kb)

Soria R., Wu K. Astron. Astrophys. 410, 53 (2003) Properties of discrete X-ray sources in the starburst spiral galaxy M83.

J/A+A/410/343/  tables.dat  (136Kb)

Vienne A., Noyelles B., Amosse A. Astron. Astrophys., 410, 343-347 (2003) Observation of 13 mutual events of Jovian satellites performed at Lille Observatory.

J/A+A/410/349/  table1.dat table4.dat table4.ps table4.tex  (627Kb)

Aggarwal K.M., Keenan F.P., Msezane A.Z. Astron. Astrophys. 410, 349 (2003) Effective collision strengths for transitions in Fe XV.

J/A+A/410/359/  table3.dat table4.dat  (34Kb)

Palmeri P., Mendoza C., Kallman T.R., Bautista M.A., Melendez M. Astron. Astrophys. 410, 359 (2003) Modeling of the iron K lines: radiative and Auger decay data for Fe II-Fe IX.

J/A+A/410/523/  table1.dat  (111Kb)

Kaltcheva N.T. Astron. Astrophys. 410, 523 (2003) The Carina Spiral Feature: Stroemgren-Hβphotometry approach. I. The photometric data-base

J/A+A/410/527/  refs.dat table3.dat table6.dat  (50Kb)

Bensby T., Feltzing S., Lundstrom I. Astron. Astrophys. 410, 527 (2003) Elemental abundance trends in the Galactic thin and thick disks as traced by nearby F and G dwarf stars

J/A+A/410/597/  refs.dat table1.dat table2.dat  (22Kb)

Walsh A.J., Macdonald G.H., Alvey N.D.S., Burton M.G., Lee J.-K. Astron. Astrophys., 410, 597-610 (2003) Observations of warm dust near methanol masers.

J/A+A/410/671/  refs.dat table1.dat table2.dat  (47Kb)

Messina S., Pizzolato N., Guinan E.F., Rodon M. Astron. Astrophys. 410, 671 (2003) Dependence of coronal X-ray emission on spot-induced brightness variations in cool main sequence stars.

J/A+A/410/795/  table2.dat  (41Kb)

Demers S., Battinelli P., Letarte B. Astron. Astrophys. 410, 795 (2003) Carbon star survey in the Local Group. VII. NGC 3109 a galaxy without a stellar halo.

J/A+A/410/813/  table4.dat  (17Kb)

Aharonian F.A., Akhperjanian A.G., Beilicke M., Bernloehr K., Boerst H.-G., Bojahr H., Bolz O., Coarasa J.A., Contreras J.L., Cortina J., Denninghoff S., Fonseca M.V., Girma M., Goebel F., Goetting N., Heinzelmann G., Hermann G., Heusler A., Hofmann W., Horns D., Jung I., Kankanyan R., Kestel M., Kettler J., Kohnle A., Konopelko A., Kranich D., Krawczynski H., Lampeitl H., Lopez M., Lorenz E., Lucarelli F., Mang O., Meyer H., Mirzoyan R., Moralejo A., Ona-Wilhelmi E., Paneque D., Panter M., Plyasheshnikov A., Puehlhofer G., de los Reyes R., Rhode W., Ripken J., Rowell G.P., Sahakian V., Samorski M., Schilling M., Schweizer T., Sevilla I., Siems M., Sobczynska D., Stamm W., Tluczykont M., Tonello N., Vitale V., Voelk H.J., Wagner R.M., Wiedner C.A.,W.Wittek. Astron. Astrophys. 410, 813 (2003) TeV gamma-ray light curve and energy spectrum of Mkn 421 during its 2001 flare as measured with HEGRA CT1.

J/A+A/410/847/  maps.dat maps/*  (71.9 Mb)

Platania P., Burigana C., Maino D., Caserini E., Bersanelli M., Cappellini B., Mennella A. Astron. Astrophys. 410, 847 (2003) Full sky study of diffuse Galactic emission at decimeter wavelengths.

J/A+A/410/879/ ngc2126.dat table3.dat var1.dat var2.dat var3.dat var4.dat var5.dat var6.dat (165Kb) Gaspar A., Kiss L.L., Bedding T.R., Derekas A., Kaspi S., Kiss C., Sarneczky K., Szabo G.M., Varadi M. Astron. Astrophys. 410, 879 (2003) The first CCD photometric study of the open cluster NGC 2126.

J/A+A/410/905/  table1.dat table2.dat  (33Kb)

Patriarchi P., Morbidelli L., Perinotto M. Astron. Astrophys. 410, 905 (2003) A study of R_{V} in galactic O stars from the 2MASS catalogue.

J/A+A/410/1039/  table1.dat table2.dat table3.dat table4.dat table5.dat table6.dat  (17Kb)

Perrier C., Sivan J.P., Naef D., Beuzit J.L., Mayor M., Queloz D., Udry S. Astron. Astrophys. 410, 1039 (2003) The ELODIE survey for northern extrasolar planets. I. 6 new extra-solar planet candidates.

J/A+A/410/1051/  table1.dat  (7Kb)

Naef D., Mayor M., Korzennik S.G., Queloz D., Udry S., Nisenson P., Noyes R.W., Brown T.M., Beuzit J.L., Perrier C., Sivan J.P. Astron. Astrophys. 410, 1051 (2003) The ELODIE survey for northern extra-solar planets. II: A Jovian planet on a long-period orbit around GJ 777A.

J/A+A/411/L59/  catalog.dat catalog.ps catalog.tex  (1.3 Mb)

Ebisawa K., Bourban G., Bodaghee A., Mowlavi N., Courvoisier T.J.-L. Astron. Astrophys. 411, L59 (2003) High-energy sources before INTEGRAL. INTEGRAL reference catalog.

J/A+A/411/361/  table1.dat  (42Kb)

Kanbur S., Ngeow C., Nikolaev S., Tanvir N.R., Hendry M.A. Astron. Astrophys. 411, 361 (2003) The extra-galactic Cepheid distance scale from LMC and Galactic period-luminosity relations.

J/A+A/411/381/  table1.dat table2.dat  (26Kb)

Sauty S., Casoli F., Boselli A., Gerin M., Lequeux J., Braine J., Gavazzi G., Dickey J., Kazes I., Fouque P. Astron. Astrophys. 411, 381 (2003) 12CO(1-0) observation of isolated late-type galaxies.

J/A+A/411/391/  notes.dat table1.dat  (78Kb)

Leon S., Verdes-Montenegro L. Astron. Astrophys. 411, 391 (2003) Revised positions for CIG galaxies.

J/A+A/411/447/  refs.dat table1.dat table2.dat  (37Kb)

Lallement R., Walsh B.Y., Vergely J.L., Crifo F., Sfeir D. Astron. Astrophys. 411, 447 (2003) 3D mapping of the dense interstellar gas around the Local Bubble

J/A+A/411/559/  calib.dat table1.dat  (31Kb)

Kovtyukh V.V., Soubiran C., Belik S.I., Gorlova N.I. Astron. Astrophys. 411, 559 (2003) High precision effective temperatures for 181 F-K dwarfs from line-depth ratios.

J/A+A/411/565/  table7.dat table7.tex  (18Kb)

Neiner C., Henrichs H.F., Floquet M., Fremat Y., Preuss O., Hubert A.M., Geers V.C., Tijani A.H., Nichols J., Jankov S. Astron. Astrophys. 411, 565 (2003) Rotation, pulsations and magnetic field in V2052 Ophiuchi: a new He-strong star.

J/A+A/412/45 See VII/237/

J/A+A/412/57 See VII/238/

J/A+A/412/399 See VII/235/

J/A+A/412/405/ (2Kb) Yudin R.V., Evans A., Barrett P., Albinson J.S., Davies J.K., Hutchinson M.G. Astron. Astrophys., 412, 405-416 (2003) Polarimetry of evolved stars. III. RV Tau and R CrB stars.

J/A+A/412/447/  table4.dat table5.dat  (25Kb)

Gerbaldi M., Faraggiana R., Lai O. Astron. Astrophys., 412, 447-464 (2003) The heterogeneous class of λBootis stars.

J/A+A/412/495/  tablea1.dat  (11Kb)

Morel T., Micela G., Favata F., Katz D., Pillitteri I. Astron. Astrophys., 412, 495-512 (2003) The photospheric abundances of active binaries. II. Atmospheric parameters and abundance patterns for 6

single-lined RS CVn systems.

J/A+A/412/633/  table1.dat  (17Kb)

Zavala J., Avila-Reese V., Hernandez-Toledo H., Firmani C. Astron. Astrophys. 412, 633 (2003) The luminous and dark matter content of disk galaxies.

J/A+A/412/657/  table1.dat table4.dat table23.dat  (17Kb)

Salome P., Combes F. Astron. Astrophys. 412, 657 (2003) Cold molecular gas in cooling flow clusters of galaxies.

J/A+A/412/707/  refs.dat table1.dat table2.txt table2z.dat  (2.7 Mb)

Ryabinkov A.I., Kaminker A.D., Varshalovich D.A. Astron. Astrophys. 412, 707 (2003) A Catalogue of absorption-line systems in QSO spectra – CoALSs.

J/A+A/412/721/  table4.dat table5.dat table6.dat table7.dat table8.dat  (87Kb)

Paunzen E., Pintado O.I., Maitzen H.M. Astron. Astrophys., 412, 721-725 (2003) CCD photometric search for peculiar stars in open clusters. V. NGC 2099, NGC 3114, NGC 6204, NGC 6705 and NGC 6756.

J/A+A/412/813/  tablea1.dat  (15Kb)

Reiners A., Schmitt J.H.M.M. Astron. Astrophys. 412, 813 (2003) Differential rotation in rapidly rotating F-stars

J/A+A/413/159/  stars.dat tablea1.dat tablea1.ps tablea1.tex  (75Kb)

Andrievsky S.M., Luck R.E., Martin P., Lepine J.R.D. Astron. Astrophys., 413, 159-172 (2004) The Galactic abundance gradient from Cepheids. V. Transition zone between 10 and 11 kpc.

J/A+A/413/233/  refs.dat table2.dat table3.dat  (18Kb)

Blaszkiewicz, L., Kus, A. J. Astron. Astrophys. 413, 233 (2004) 12.2 GHz survey towards 6.7 GHz methanol masers. A comparison of 12.2 GHz and 6.7 GHz spectra.

J/A+A/413/285/  table2.dat  (100Kb)

Gulliver A.F., Adelman S.J., Friesen T.P. Astron. Astrophys. 413, 285 (2004) A spectroscopic atlas of o Pegasi (A1 IV) λλ3826-4882.

J/A+A/413/293/  table3.dat  (87Kb)

Lehmann H., Mkrtichian D.E. Astron. Astrophys., 413, 293-299 (2004) The eclipsing binary star RZ Cas. I. First spectroscopic detection of rapid pulsations in an Algol system.

J/A+A/413/343/  table1.dat table2.dat  (25Kb)

Cacciari C., Bragaglia A., Rossetti E., Fusi Pecci F., Mulas G., Carretta E., Gratton R.G., Momany Y., Pasquini L. Astron. Astrophys., 413, 343-362 (2004) Mass motions and chromospheres of RGB stars in the globular cluster NGC 2808.

J/A+A/413/771/  table.dat  (13Kb)

Metz M., Geffert M. Astron. Astrophys. 413, 771 (2004) Formalism and quality of a proper motion link with extragalactic objects for astrometric satellite missions.

J/A+A/413/779/ table3.txt table3a.dat table3b.dat table4.dat table4.txt table5.dat table5.txt table7.dat table8a.dat table8b.dat table78.txt (58.9 Mb) Nahar S.N. Astron. Astrophys., 413, 779-787 (2004) Atomic data from the Iron Project. LIV. Relativistic calculations for allowed and forbidden fine structure transitions in Fe XX.

J/A+A/413/843/  table4.dat table5.dat table6.dat table7.dat table8.dat table9.dat  (21Kb)

Kawara K., Matsuhara H., Okuda H., Taniguchi Y., Sato Y., Sofue Y., Wakamatsu K., Oyabu S., Sanders D.B., Cowie L.L. Astron. Astrophys., 413, 843-859 (2004) ISO deep far-infrared survey in the ”Lockman Hole”. III. Catalogs and source counts at 90 + 170μm.

J/A+A/413/1037/  table1.dat table2.dat  (184Kb)

Kimeswenger S., Lederle C., Richichi A., Percheron I., Paresce F., Armsdorfer B., Bacher A., Cabrera-Lavers A.L., Kausch W., Rassia E., Schmeja S., Tapken C., Fouque P., Maury A., Epchtein N. Astron. Astrophys. 413, 1037 (2004) J-K DENIS photometry of a VLTI-selected sample of bright southern stars.

J/A+A/413/1177/  table4.dat table5.dat table6.dat table7.dat table8.dat tables.ps tables.tex  (73Kb)

Dore L., Caselli P., Beninati S., Bourke T., Myers P.C., Cazzoli G. Astron. Astrophys. 413, 1177 (2004) Laboratory and radio-astronomical spectroscopy of the hyperfine structure of N_{2}D^{+}.

J/A+A/414/23/  table3.dat  (43Kb)

James P.A., Shane N.S., Beckman J.E., Cardwell A., Collins C.A., Etherton J., de Jong R.S., Fathi K., Knapen J.H., Peletier R.F., Percival S.M., Pollacco D.L., Seigar M.S., Stedman S., Steele I.A Astron. Astrophys. 414, 23 (2004) The HαGalaxy Survey. I. The galaxy sample, Hαnarrow-band observations and star formation parameters for 334 galaxies

J/A+A/414/235/  table1.dat table2.dat table3.dat table4.dat  (23Kb)

Szymczak M., Gerard E. Astron. Astrophys. 414, 235 (2004) Observations of hydroxyl ground state transitions in a complete sample of methanol sources.

J/A+A/414/351/ table1.dat table2.dat table3.dat table4.dat table5.dat table6.dat table7.dat table8.dat table9.dat (35Kb) Naef D., Mayor M., Beuzit J.L., Perrier C., Queloz D., Sivan J.P., Udry S. Astron. Astrophys., 414, 351-359 (2004) The ELODIE survey for northern extra-solar planets. III. Three planetary candidates detected with ELODIE.

J/A+A/414/377/  table4.dat table5.dat  (49Kb)

Mendoza C., Kallman T.R., Bautista M.A., Palmeri P. Astron. Astrophys. 414, 377 (2004) Decay properties of K-vacancy states in Fe X-Fe XVII

J/A+A/414/487/  table2.dat table3.dat table4.dat  (38Kb)

Veron-cetty M.-P., Balayan S.K., Mickaelian A.M., Mujica R., Chavushyan V., Hakopian S.A., Engels D., Veron P., Zickgraf F.-J., Voges W., Xu D.-W. Astron. Astrophys., 414, 487-496 (2004) Optically bright active galactic nuclei in the ROSAT-Faint Source Catalogue.

J/A+A/414/503/  table1.dat table3.dat  (15Kb)

Bonifacio P., Sbordone L., Marconi G., Pasquini L., Hill V. Astron. Astrophys. 414, 503 (2004) The Sgr dSph hosts a metal-rich population.

J/A+A/414/699/  hrdat/* hrfit/* lrdat/* lrfit/* table2.dat  (9.5 Mb)

Cincunegui C., Mauas P.J.D. Astron. Astrophys. 414, 699 (2004) Library of flux-calibrated echelle spectra of southern late-type dwarfs with different activity levels.

J/A+A/414/825/  table1.dat table2.dat  (16Kb)

Satyapal S., Sambruna R.M., Dudik R.P. Astron. Astrophys. 414, 825 (2004) A joint mid-infrared spectroscopic and X-ray imaging investigation of LINER galaxies.

J/A+A/414/905/  table1.dat  (20Kb)

Hunt L.K., Pierini D., Giovanardi C. Astron. Astrophys., 414, 905-918 (2004) Near-infrared observations of galaxies in Pisces-Perseus. V. On the origin of bulges.

J/A+A/414/949/  refs.dat table1.dat table2.dat  (14Kb)

Weselak T., Galazutdinov G.S., Musaev F.A., Krelowski J. Astron. Astrophys. 414, 949 (2004) On the relation between diffuse bands and column densities of H2, CH and CO molecules.

J/A+A/415/9/  notes.dat table1.dat table2.dat  (114Kb)

Woudt P.A., Kraan-Korteweg R.C., Cayatte V., Balkowski C., Felenbok P. Astron. Astrophys. 415, 9 (2004) Extragalactic Large-Scale Structures behind the Southern Milky Way. IV. Redshifts Obtained with MEFOS.

J/A+A/415/63/  table2.dat table3.dat table4.dat tables.tex  (44Kb)

Mollenhoff C. Astron. Astrophys. 415, 63 (2004) Disk-bulge decompositions of spiral galaxies in UBVRI.

J/A+A/415/123/  booterr.dat index.dat phot.dat syserr.dat  (127Kb)

Puzia T.H., Kissler-Patig M., Thomas D., Maraston C., Saglia R.P., Bender R., Richtler T., Goudfrooij P., Hempel M. Astron. Astrophys. 415, 123 (2004) VLT spectroscopy of globular cluster systems. I. The Photometric and Spectroscopic Data Set.

J/A+A/415/145/  stars.dat table4.dat  (15Kb)

Boeche C., Munari U., Tomasella L., Barbon R. Astron. Astrophys. 415, 145 (2004) Kinematics and binaries in young stellar aggregates. II. NGC 6913 = M29.

J/A+A/415/155/  table4.dat  (15Kb)

Bensby T., Feltzing S., Lundstrom I. Astron. Astrophys. 415, 155 (2004) Oxygen trends in the Galactic thin and thick disks.

J/A+A/415/241/  table1.dat  (104Kb)

Aerts C., Waelkens C., Daszynska-Daszkiewicz J., Dupret M.-A., Thoul A., Scuflaire R., Uytterhoeven K., Niemczura E., Noels A. Astron. Astrophys. 415, 241 (2004) Asteroseismology of the βCep star HD 129929. I. Observations, oscillation frequencies and stellar parameters.

J/A+A/415/273/  table2.dat table3.dat table4.dat  (21Kb)

Leedjarv L., Burmeister M., Mikolajewski M., Puss A., Annuk K., Galan C. Astron. Astrophys. 415, 273 (2004) Emission lines in the spectrum of the symbiotic star AG Draconis from 1997 to 2003

J/A+A/415/283/  observ.dat  (11Kb)

Djurasevic G., Albayrak B., Tanriverdi T., Erkapic S. Astron. Astrophys. 415, 283 (2004) A Photometric Study of NN Virginis.

J/A+A/415/289/  table4.dat  (145Kb)

Harmanec P., Bozic H., Thanjavur K., Robb R.M., Ruzdjak D., Sudar D. Astron. Astrophys. 415, 289 (2004) An improved ephemeris and physical elements of ER Vul.

J/A+A/415/331/  table4.dat  (75Kb)

Pagano I., Linsky J.L., Valenti J., Duncan D.K. Astron. Astrophys. 415, 331 (2004) HST/STIS High Resolution Echelle Spectra of alpha Centauri A (G2 V).

J/A+A/415/391/ stars.dat table1.dat table2.dat table3.dat table4.dat table5.dat table6.dat table7.dat table8.dat table9.dat table10.dat table11.dat table12.dat table13.dat (45Kb) Mayor M., Udry S., Naef D., Pepe F., Queloz D., Santos N.C., Burnet M. Astron. Astrophys., 415, 391-402 (2004) The CORALIE survey for southern extra-solar planets. XII. Orbital solutions for 16 extra-solar planets discovered with CORALIE.

J/A+A/415/521/ table1.dat table10.dat table11.dat table12.dat table13.dat table14.dat table15.dat table16.dat table17.dat table18.dat table19.dat table20.dat table21.dat table22.dat table23.dat table24.dat table25.dat table26.dat table27.dat (69Kb) Storm J., Carney B.W., Gieren W.P., Fouque P., Freedman W.L., Madore B.F., Habgood M.J. Astron. Astrophys., 415, 521-529 (2004) BVRIJK light curves and radial velocity curves for selected Magellanic Cloud Cepheids.

J/A+A/415/531/ stars.dat tablea2.dat tablea3.dat tablea4.dat tablea5.dat tablea6.dat tablea7.dat tablea8.dat tablea9.dat tablea10.dat tablea11.dat (33Kb) Storm J., Carney B.W., Gieren W.P., Fouque P., Latham D.W., Fry A.M. Astron. Astrophys., 415, 531-547 (2004) The effect of metallicity on the Cepheid Period-Luminosity relation from a Baade-Wesselink analysis of Cepheids in the Galaxy and in the Small Magellanic Cloud.

J/A+A/415/549/  table5.dat table6.dat  (23Kb)

Ricci R., Prandoni I., Gruppioni C., Sault R.J., De Zotti G. Astron. Astrophys., 415, 549-558 (2004) High-frequency polarization properties of southern Kuehr sources.

J/A+A/415/849/  refs.dat table7a.dat table7b.dat  (34Kb)

Barbera C., Athanassoula E., Garcia-Gomez C. Astron. Astrophys. 415, 849 (2004) Deprojecting spiral galaxies using Fourier analysis. Application to the Frei sample.

J/A+A/415/863/  comp/* flux/* mag/* table1.dat table3.dat  (200Kb)

Germany L.M., Reiss D.J., Schmidt B.P., Stubbs C.W., Suntzeff N.B. Astron. Astrophys. 415, 863 (2004) Results from the Mount Stromlo Abell Cluster Supernova Search.

J/A+A/415/993/  stars.dat tablea1.dat  (14Kb)

Nissen P.E., Chen Y.Q., Asplund M., Pettini M. Astron. Astrophys. 415, 993 (2004) Sulphur and zinc abundances in Galactic stars and damped Ly-alpha systems.

J/A+A/415/1153/  table2.dat table3.dat table4.dat table5.dat  (24Kb)

Santos N.C., Israelian G., Mayor M. Astron. Astrophys. 415, 1153 (2004) Spectroscopic [Fe/H] for 98 extra-solar planet-host stars. Exploring the probability of planet formation.

J/A+A/416/35/  table3.dat table4.dat  (195Kb)

Wadadekar Y. Astron. Astrophys. 416, 35 (2004) Search for radio emission from AGN in the FIRST Survey Area

J/A+A/416/111/  table2.dat  (17Kb)

Battinelli P., Demers S. Astron. Astrophys., 416, 111-118 (2004) Carbon star survey of Local Group galaxies. X. Wolf-Lundmark-Melotte a galaxy with an extreme C/M ratio.

J/A+A/416/125/  tableb1.dat tableb2.dat tableb3.dat tableb4.dat  (98Kb)

Deacon N.R., Hambly N.C. Astron. Astrophys., 416, 125-136 (2004) Proper motion surveys of the young open clusters Alpha Persei and the Pleiades.

J/A+A/416/281/  table4.dat table5.dat table6.dat table7.dat tables.ps tables.tex  (155Kb)

Sanz-forcada J., Favata F., Micela G. Astron. Astrophys., 416, 281-290 (2004) Coronal versus photospheric abundances of stars with different activity levels.

J/A+A/416/383/  table5.dat table6.dat table7.dat table8.dat  (665Kb)

Jonauskas V., Keenan F.P., Foord M.E., Heeter R.F., Rose S.J., van Hoof P.A.M., Ferland G.J., Aggarwal K.M., Kisielius R., Norrington P.H. Astron. Astrophys. 416, 383 (2004) Relativistic allowed and forbidden transition probabilities for fluorine-like Fe XVIII.

J/A+A/416/475/  table5.dat  (23Kb)

Marquez I., Durret F., Masegosa J., Moles M., Varela J., Gonzalez Delgado R.M., Maza J., Perez E., Roth M. Astron. Astrophys. 416, 475 (2004) Long slit spectroscopy of a sample of isolated spirals with and without an AGN.

J/A+A/416/507/  table5.dat  (29Kb)

Coccato L., Corsini E.M., Pizzella A., Morelli L., Funes, J.G., Bertola F. Astron. Astrophys. 416, 507 (2004) Minor-axis velocity gradients in disk galaxies.

J/A+A/416/515/  table4.dat table5.dat table6.dat  (36Kb)

Davoust E., Contini T. Astron. Astrophys. 416, 515 (2004) Starbursts in barred spiral galaxies. VI. HI observations and the K-band Tully-Fisher relation.

J/A+A/416/537/  table3.dat table4.dat table5.dat table7.dat  (400Kb)

Larsen S.S. Astron. Astrophys. 416, 537 (2004) The structure and environment of young stellar clusters in spiral galaxies.

J/A+A/416/555/  table4.dat table6.dat  (20Kb)

Lopez-Marti B., Eisloeffel J., Scholz A., Mundt R. Astron. Astrophys., 416, 555-576 (2004) The brown dwarf population in the Chamaeleon I cloud.

J/A+A/416/655/  table1.dat table2.dat tables.ps tables.tex  (142Kb)

Lyubchik Y., Jones H.R.A., Pavlenko Y.V, Viti S., Pickering J.C., Blackwell-Whitehead R. Astron. Astrophys. 416, 655 (2004) Atomic lines in infrared spectra for ultracool dwarfs

J/A+A/416/669/  table2.dat table4a.dat table4b.dat  (87Kb)

Bozic H., Harmanec P., Yang S., Ziznovsky J., Percy J., Ruzdjak D., Sudar D., Slechta M., Skoda P., Krpata J., Buil C. Astron. Astrophys. 416, 669 (2004) Properties and nature of Be stars. XXII. Long-term light and spectral variations of the new bright Be star HD 6226.

J/A+A/416/811/  table1.dat table2.dat  (33Kb)

Abad C., Docobo J.A., Lanchares V., Lahulla J.F., Abelleira P., Blanco J., Alvarez C. Astron. Astrophys. 416, 811 (2004) Reduction of CCD Observations of Visual Binaries, using the ”Tepui” Function as PSF.

J/A+A/416/839/  table1.dat  (14Kb)

Boschin W., Girardi M., Barrena R., Biviano A., Feretti L., Ramella M. Astron. Astrophys. 416, 839 (2004) Internal dynamics of the radio-halo cluster A2219: a multiwavelength analysis.

J/A+A/416/917/  refs.dat table2.dat table3.dat table4.dat  (286Kb)

Galleti S., Federici L., Bellazzini M., Fusi Pecci F., Macrina S. Astron. Astrophys. 416, 917 (2004) 2MASS NIR photometry for 693 candidate globular clusters in M31 and the Revised Bologna Catalogue.

J/A+A/416/1037/  table3.dat  (11Kb)

Hulleman F., van Kerkwijk M.H., Kulkarni S.R. Astron. Astrophys. 416, 1037 (2004) The anomalous X-ray pulsar 4U 0142+61: variability in the infrared and a spectral break in the optical.

J/A+A/416/1117/  table2.dat table3.dat table8.dat  (130Kb)

Cayrel R., Depagne E., Spite M., Hill V., Spite F., Francois P., Plez B., Beers T., Primas F., Andersen J., Barbuy B., Bonifacio P., Molaro P., Nordstrom B. Astron. Astrophys. 416, 1117 (2004) First stars. V. Abundance patterns from C to Zn and supernova yields in the early Galaxy.

J/A+A/417/107/  table2.dat table3.dat  (89Kb)

Russeil D., Castets A. Astron. Astrophys. 417, 107 (2004) CO observations in direction of southern HII regions.

J/A+A/417/115/  table2.dat  (16Kb)

Williams S.J., Fuller G.A., Sridharan T.K. Astron. Astrophys. 417, 115 (2004) The circumstellar environments of HMPOs. I: Submillimetre continuum emission.

J/A+A/417/263/ notes.dat refs.dat table1.dat table2.dat table3.dat table4.dat table5.dat table6a.dat table6b.dat table6c.dat table6d.dat (766Kb) Budding E., Erdem A., Cicek C., Bulut I., Soydugan F., Soydugan E., Bakis V., Demircan O. Astron. Astrophys., 417, 263-268 (2004) Catalogue of Algol type binary stars.

J/A+A/417/269/  table4.dat  (22Kb)

Reyniers M., Van Winckel H., Gallino R., Straniero O. Astron. Astrophys. 417, 269 (2004) A study of the s-process in the carbon-rich post-AGB stars IRAS 06530-0213 and IRAS 08143-4406 on the basis of VLT-UVES spectra.

J/A+A/417/375/  table1.dat table2.dat  (49Kb)

Milovanovic N., Dimitrijevic M.S., Popovic L.C., Simic Z. Astron. Astrophys. 417, 375 (2004) Importance of collisions with charged particles for stellar UV line shapes: Cd III.

J/A+A/417/421/  table2.dat  (14Kb)

Braun R., Thilker D.A. Astron. Astrophys., 417, 421-435 (2004) The WSRT wide-field HI survey. II. Local Group features.

J/A+A/417/479/  table2.dat  (17Kb)

Battinelli P., Demers S.

Astron. Astrophys. 417, 479 (2004) Carbon star survey of Local Group galaxies. IX. The spheroidal galaxy NGC 185.

J/A+A/417/499/  spectra/* table6.dat table7.dat table8.dat table9.dat  (6.0 Mb)

Gavazzi G., Zaccardo A., Sanvito G., Bonfanti C., Boselli A. Astron. Astrophys. 417, 499 (2004) Spectrophotometry of galaxies in the Virgo cluster. II: The data.

J/A+A/417/515/  iwz1.fit  (24Kb)

Veron-Cetty M.-P., Joly M., Veron P. Astron. Astrophys. 417, 515 (2004) The unusual emission line spectrum of I Zw 1.

J/A+A/417/557/  table4.dat table6.dat  (1.7 Mb)

Lamm M.H., Bailer-Jones C.A.L., Mundt R., Herbst W., Scholz A. Astron. Astrophys. 417, 557 (2004) A Rotational and Variability Study of a Large Sample of PMS Stars in NGC 2264

J/A+A/417/597/ (2Kb) Recio-Blanco A., Piotto G., Aparicio A., Renzini A. Astron. Astrophys., 417, 597-604 (2004) Rotation velocities of hot horizontal branch stars in the globular clusters NGC 1904, NGC 2808, NGC 6093, and NGC 7078: The database.

J/A+A/417/615/  table3.dat  (10Kb)

Codella C., Lorenzani A., Gallego A.T., Cesaroni R., Moscadelli L. Astron. Astrophys., 417, 615-624 (2004) The association between masers and outflows in massive star forming regions.

J/A+A/417/651/  table1.dat table2.dat table3.dat  (31Kb)

Schmitt J.H.M.M., Liefke C. Astron. Astrophys., 417, 651-665 (2004) NEXXUS: A comprehensive ROSAT survey of coronal X-ray emission among nearby solar-like stars.

J/A+A/417/745/  imag.dat rmag.dat vmag.dat  (35Kb)

Csizmadia Sz., Patkos L., Moor A., Konyves V. Astron. Astrophys. 417, 745 (2004) Photometric study of the contact binary star V861 Herculis.

J/A+A/417/827/  di.dat  (117Kb)

Itoh N., Nozawa S. Astron. Astrophys., 417, 827-832 (2004) Relativistic corrections to the Sunyaev-Zeldovich effect for extremely hot clusters of galaxies.

J/A+A/417/945/  table5.dat table6.dat  (262Kb)

Prisinzano L., Micela G., Sciortino S., Favata F. Astron. Astrophys., 417, 945-960 (2004) Parameter properties and stellar population of the old open cluster NGC 3960.

J/A+A/417/961/  table3a.dat table3b.dat  (833Kb)

Baume G., Moitinho A., Giorgi E.E., Carraro G., Vazquez R.A. Astron. Astrophys., 417, 961-972 (2004) NGC 2580 and NGC 2588. Two open clusters in the Third Galactic Quadrant.

J/A+A/418/1/  table3.dat  (13Kb)

Vollmer B., Thierbach M., Wielebinski R. Astron. Astrophys. 418, 1 (2004) Radio continuum spectra of galaxies in the Virgo cluster region.

J/A+A/418/33/  table2.dat  (27Kb)

Battinelli P., Demers S. Astron. Astrophys. 418, 33 (2004) Carbon star survey in the Local Group. VIII. Probing the stellar halo of NGC 147.

J/A+A/418/103/  table2.dat  (61Kb)

Messineo M., Habing H.J., Menten K.M., Omont A., Sjouwerman L.O. Astron. Astrophys. 418, 103 (2004) 86 GHz SiO maser survey of late-type stars in the Inner Galaxy. II. Infrared photometry of the SiO target stars

J/A+A/418/131/  refs.dat table2n.dat tablea.dat tablec.dat  (156Kb)

Kiss C., Moor A., Toth L.V.

Astron. Astrophys. 418, 131 (2004) Far-infrared loops in the 2nd Galactic Quadrant.

J/A+A/418/249/  table2.dat  (48Kb)

Arentoft T., Lampens P., Van Cauteren P., Duerbeck H.W., Garcia-Melendo E., Sterken C Astron. Astrophys. 418, 249 (2004) On the Delta Scuti star in the eclipsing binary WX Eridani.

J/A+A/418/357 See I/290/

J/A+A/418/363/  table2.dat table2.ps table2.tex  (173Kb)

Tayal S.S. Astron. Astrophys. 418, 363 (2004) Electron-impact excitation of fine-structure levels in Cl II.

J/A+A/418/371/  table1.dat table2.dat table3.dat table4.dat tables.ps tables.tex  (286Kb)

Aggarwal K.M., Keenan F.P. Astron. Astrophys. 418, 371 (2004) Electron impact excitation of Fe XIII.

J/A+A/418/445/  fcos2.dat  (17Kb)

Mieske S., Hilker M., Infante L. Astron. Astrophys., 418, 445-458 (2004) Fornax compact object survey FCOS: On the nature of Ultra Compact Dwarf galaxies

J/A+A/418/465/  refs.dat table1.dat  (13Kb)

Lutz D., Maiolino R., Spoon H.W.W., Moorwood A.F.M. Astron. Astrophys. 418, 465 (2004) The relation between AGN hard X-ray emission and mid-infrared continuum from ISO spectra: Scatter and unification aspects.

J/A+A/418/509/  notes.dat table1.dat table4.dat  (29Kb)

Van Den Berg M., Tagliaferri G., Belloni T., Verbunt F. Astron. Astrophys., 418, 509-523 (2004) A Chandra observation of the old open cluster M67.

J/A+A/418/525/  table1.dat table2.dat  (7Kb)

Carraro G., Romaniello M., Ventura P., Patat F. Astron. Astrophys., 418, 525-537 (2004) The star cluster Collinder 232 in the Carina complex and its relation to Trumpler 14/16.

J/A+A/418/551/  table3.dat table8.dat table8.tex  (60Kb)

Mishenina T.V., Soubiran C., Kovtyukh V.V., Korotin S.A. Astron. Astrophys., 418, 551-562 (2004) On the correlation of elemental abundances with kinematics among Galactic disk stars.

J/A+A/418/841/  table4.dat  (827Kb)

Naze Y., Antokhin I.I., Rauw G., Chu Y.-H., Gosset E., Vreux J.-M. Astron. Astrophys. 418, 841 (2004) XMM-Newton observations of the giant H II region N11 in the LMC.

J/A+A/418/869/  table3.dat  (113Kb)

Boschi F., Munari U. Astron. Astrophys. 418, 869 (2004) M31-RV evolution and its alleged multi-outburst pattern.

J/A+A/418/885/  fdfss/* table1.dat  (8.8 Mb)

Noll S., Mehlert D., Appenzeller I., Bender R., Boehm A., Gabasch A., Heidt J., Hopp U., Jaeger K., Seitz S., Stahl O., Tapken C., Ziegler B.L. Astron. Astrophys. 418, 885 (2004) The FORS Deep Field Spectroscopic Survey

J/A+A/418/927/  table2.dat  (49Kb)

Revnivtsev M., Sazonov S., Jahoda K., Gilfanov M. Astron. Astrophys. 418, 927 (2004) RXTE All-Sky Slew Survey. Catalog of X-ray sources at —b— 10^{∘}.

J/A+A/418/979/  table1.dat table2.dat  (248Kb)

Piatti A.E., Claria J.J., Ahumada A.V. Astron. Astrophys. 418, 979 (2004) The relatively young, metal-poor and distant open cluster NGC 2324.

J/A+A/418/989 See V/117/

J/A+A/418/1073/  table2.dat table3.dat  (89Kb)

Wahlgren G.M., Hubrig S. Astron. Astrophys. 418, 1073 (2004) Emission lines in the optical spectrum of 3 Cen A.

J/A+A/418/1171/ elevels.dat table1b.dat table2b.dat table3b.dat table4b.dat table5b.dat table6b.dat (369Kb) Bautista M., Mendoza C., Kallman T.R., Palmeri P. Astron. Astrophys. 418, 1171 (2004) K-shell photoionization and electron impact excitation of Fe XVIII-Fe XXIII.

J/A+A/419/47/  table1.dat table2.dat table3.dat table4.dat  (23Kb)

Finoguenov A., Briel U.G., Henry J.P., Gavazzi G., Iglesias-Paramo J., Boselli A. Astron. Astrophys., 419, 47-61 (2004) The X-ray luminosity function of galaxies in the Coma cluster.

J/A+A/419/71/  notes.dat table3.dat table4.dat  (27Kb)

Giacintucci S., Venturi T., Bardelli S., Dallacasa D., Zucca E. Astron. Astrophys. 419, 71 (2004) AGN and starburst radio activity in the A3558 cluster complex

J/A+A/419/139/  m3.dat m5.dat m10.dat m13.dat m92.dat  (326Kb)

Valenti E., Ferraro F.R., Perina S., Origlia L. Astron. Astrophys. 419, 139 (2004) Near-infrared photometry of five Galactic globular clusters.

J/A+A/419/149/  table2.dat  (453Kb)

Villanova S., Baume G., Carraro G., Geminale A. Astron. Astrophys., 419, 149-159 (2004) Optical photometry and spectral classification in the field of the open cluster NGC 6996 in the North America Nebula.

J/A+A/419/167/  notes.dat table1.dat  (130Kb)

Zapatero Osorio M.R., Martin E.L. Astron. Astrophys. 419, 167 (2004) A CCD imaging search for wide metal-poor binaries.

J/A+A/419/215/  table2.dat table3.dat table23.ps  (392Kb)

Hartman H., Gull T., Johansson S., Smith N., HST Eta Carinae Treasury Project Team Astron. Astrophys. 419, 215 (2004) Identification of emission lines in the low-ionization strontium filament near eta Carinae.

J/A+A/419/249/  table7.dat  (16Kb)

Scholz A., Eisloeffel J. Astron. Astrophys. 419, 249 (2004) Rotation and accretion of very low mass objects in the sigma Ori cluster.

J/A+A/419/469/  table2.dat table3.dat table4.dat  (14Kb)

Pustilnik S., Kniazev A., Pramskij A., Izotov Y., Foltz C., Brosch N., Martin J.-M., Ugryumov A. Astron. Astrophys., 419, 469-484 (2004) HS 0837+4717 - a metal-deficient blue compact galaxy with large nitrogen excess.

J/A+A/419/533/  table2.dat table3.dat  (14Kb)

Blecha A., Meylan G., North P., Royer F. Astron. Astrophys. 419, 533 (2004) Palomar 13: a velocity dispersion inflated by binaries ?

J/A+A/419/563/  table1.dat table2.dat table3.dat  (18Kb)

Jacoby G.H., Van De Steene G. Astron. Astrophys., 419, 563-582 (2004) Planetary nebulae near the Galactic Center: Identifications.

J/A+A/419/607/  table2.dat table3.dat table4.dat  (33Kb)

Saad S.M., Kubat J., Koubsky P., Harmanec P., Skoda P., Korcakova D., Krticka J., Slechta M., Bozic H., Ak H., Hadrava P., Votruba V. Astron. Astrophys. 419, 607 (2004) Properties and nature of Be stars. XXIII. Long-term variations and physical properties of kappa Dra.

J/A+A/419/999/  object.dat table1.dat table2.dat table3.dat  (79Kb)

Giannini T., McCoey C., Caratti o Garatti A., Nisini B., Lorenzetti D., Flower D. Astron. Astrophys., 419, 999 (2004)

On the excitation of the infrared knots along protostellar jets.

J/A+A/419/1123/  table.dat table.ps table.tex  (40Kb)

Mooney C.J., Rolleston W.R.J., Keenan F.P., Dufton P.L., Smoker J.V., Ryans R.S.I., Aller L.H., Trundle C. Astron. Astrophys. 419, 1123 (2004) High-resolution spectroscopy of globular cluster post-asymptotic giant branch stars.

J/A+A/420/61/  table2.dat table3.dat  (20Kb)

Schuecker P., Boehringer H., Voges W. Astron. Astrophys., 420, 61-74 (2004) Detection of X-ray clusters of galaxies by matching RASS photons and SDSS galaxies within GAVO.

J/A+A/420/97/  table1.dat  (14Kb)

Boehm A., Ziegler B.L., Saglia R.P., Bender R., Fricke K.J., Gabasch A., Heidt J., Mehlert D., Noll S., Seitz S. Astron. Astrophys. 420, 97 (2004) The Tully-Fisher relation at intermediate redshift.

J/A+A/420/135/  table4.dat table4.tex  (15Kb)

Tamura T., Kaastra J.S., Den Herder J.W.A., Bleeker J.A.M., Peterson J.R. Astron. Astrophys., 420, 135-146 (2004) Elemental abundances in the intracluster medium as observed with XMM-Newton.

J/A+A/420/147/  tablea1.dat tablea2.dat tablea3.dat tablea4.dat tablea5.dat tablea6.dat  (27Kb)

Blais-Ouellette S., Amram P., Carignan C., Swaters R. Astron. Astrophys., 420, 147-161 (2004) Accurate determination of the mass distribution in spiral galaxies. III. Fabry-Perot imaging spectroscopy of 6 spiral galaxies.

J/A+A/420/173/  table2.dat  (4.7 Mb)

Sollima A., Ferraro F.R., Origlia L., Pancino E., Bellazzini M. Astron. Astrophys. 420, 173 (2004) IR photometric properties of red giants in omega Cen.

J/A+A/420/183/  sp/* table3.dat table4.dat table5.dat  (554.6 Mb)

Allende Prieto C., Barklem P.S., Lambert D.L., Cunha K. Astron. Astrophys. 420, 183 (2004) S4N: A Spectroscopic Survey of Stars in the Solar Neighborhood.

J/A+A/420/447/  table4.dat table5.dat table6.dat table7.dat  (20Kb)

Vakulik V.G., Schild R., Dudinov V.N., Minakov A.A., Nuritdinov S.N., Tsvetkova V.S., Zheleznyak A.P., Konichek V.V., Sinelnikov I.Ye., Burkhonov O.A.,B.P.Artamonov, Bruevich V.V. Astron. Astrophys. 420, 447 (2004) Color effects associated with the 1999 microlensing brightness peaks in gravitationally lensed quasar Q2237+0305.

J/A+A/420/571/  table1.dat table2.dat table3.dat table4.dat table5.dat  (19Kb)

Cassatella A., Lamers H.J.G.L.M., Rossi C., Altamore A., Gonzalez-Riestra R. Astron. Astrophys., 420, 571-588 (2004) A study of the expanding envelope of Nova V1974 Cyg 1992 based on IUE high resolution spectroscopy.

J/A+A/420/595/ (2Kb) Afsar M., Heckert P.A., Ibanoglu C. Astron. Astrophys., 420, 595-604 (2004) Long-term luminosity variations and period changes in CG Cygni.

J/A+A/420/673/  table2.dat table2.ps  (58Kb)

Nishimura M., Sadakane K., Kato K., Takeda Y., Mathys G. Astron. Astrophys. 420, 673 (2004) Chemical composition of the magnetic B star HR 5049.

J/A+A/420/683/ bd231978.dat hd008009.dat hd008071.dat hd009911.dat hd013357.dat hd017159.dat hd019440.dat hd030101.dat hd033334.dat hd066491.dat hd106515.dat hd108574.dat hd123963.dat hd132563.dat hd132844.dat hd135101.dat hd190042.dat hd200466.dat hd213013.dat hd215812.dat hd216122.dat hd219542.dat hi104687.dat stars.dat (153Kb) Desidera S., Gratton R.G., Scuderi S., Claudi R.U., Cosentino R., Barbieri M., Bonanno G., Carretta E., Endl M., Lucatello S., Martinez Fiorenzano A.F., Marzari F. Astron. Astrophys. 420, 683 (2004) Abundance difference between components of wide binaries.

J/A+A/420/763/  level.dat table.dat  (375Kb)

Bautista M.A. Astron. Astrophys., 420, 763-768 (2004) Atomic data from the IRON Project. LV. Electron impact excitation of NiII.

J/A+A/420/783/  table1.dat table2.dat table3.dat table4.dat table5.dat tables.ps tables.tex  (3.2 Mb)

Aggarwal K.M., Keenan F.P., Kisielius R. Astron. Astrophys. 420, 783 (2004) Radiative rates for transitions in Fe XVII.

J/A+A/420/873/  table1.dat table2.dat  (38Kb)

Varela J., Moles M., Marquez I., Galletta G., Masegosa J., Bettoni D. Astron. Astrophys., 420, 873-879 (2004) Properties of isolated disk galaxies.

J/A+A/420/945/  table2.dat table2.ps table2.tex  (65Kb)

Galfalk M., Olofsson G., Kaas A.A., Olofsson S., Bontemps S., Nordh L., Abergel A., Andre P., Boulanger F., Burgdorf M., Casali M.M., Cesarsky C.J., Davies J., Falgarone E., Montmerle T., Perault M., Persi P., Prusti T., Puget J.L., Sibille F. Astron. Astrophys. 420, 945 (2004) ISOCAM observations of the L1551 star formation region.

J/A+A/420/1039/  obs2001.dat obs2002.dat obs2003.dat  (47Kb)

Albayrak B., Djurasevic G., Erkapic S., Tanriverdi T. Astron. Astrophys. 420, 1039 (2004) Modeling the Changing Spot Features of SW Lacertae: A Three Year Study.

J/A+A/421/41/  tables.dat  (7Kb)

Gabasch A., Bender R., Seitz S., Hopp U., Saglia R.P., Feulner G., Snigula J., Drory N., Appenzeller I., Heidt J., Mehlert D., Noll S., Boehm A., Jaeger K., Ziegler B., Fricke K.J. Astron. Astrophys. 421, 41 (2004) The evolution of the luminosity functions in the FORS Deep Field from low to high redshift. I. The blue bands.

J/A+A/421/59/  tablea1.dat  (13Kb)

Gerken B., Ziegler B.L., Balogh M.L., Gilbank D.G., Fritz A., Jaeger K. Astron. Astrophys. 421, 59 (2004) Star formation activity of intermediate redshift cluster galaxies out to the infall regions.

J/A+A/421/175/  tablea1.dat tableb1.dat  (17Kb)

Pillitteri I., Micela G., Sciortino S., Damiani F., Harnden F.R.JR Astron. Astrophys., 421, 175-186 (2004) XMM-Newton observations of the young open cluster Blanco 1. I. X-ray spectroscopy and photometry.

J/A+A/421/455/  refs.dat table2.dat  (21Kb)

Kopylov A.I., Artyukh V.S., Kopylova F.G. Astron. Astrophys., 421, 455-459 (2004) Investigation of radio sources from the interplanetary scintillation Pushchino Survey. II. Optical identification.

J/A+A/421/509/  table3.dat  (106Kb)

Ansari R., Auriere M., Baillon P., Bouquet A., Coupinot G., Coutures C., Ghesquiere C., Giraud-Heraud Y., Gillieron D., Gondolo P., Hecquet J., Kaplan J., Kim A., Le Du Y., Melchior A.L., Moniez M., Picat J.P., Soucail G. Astron. Astrophys. 421, 509 (2004) Variable stars towards the bulge of M31: the AGAPE catalogue.

J/A+A/421/539/  addit.dat table2.dat table3.dat table4.dat table7.dat table8.dat  (32Kb)

Izotov Y.I., Papaderos P., Guseva N.G., Fricke K.J., Thuan T.X. Astron. Astrophys. 421, 539 (2004) Deep VLT spectroscopy of the blue compact dwarf galaxies Tol 1214-277 and Tol 65

J/A+A/421/595/  refs.dat table1a.dat table1b.dat  (49Kb)

Garcia-gomez C., Barbera C., Athanassoula E., Bosma A., Whyte L. Astron. Astrophys., 421, 595-601 (2004) Deprojecting spiral galaxies using Fourier analysis. Application to the Ohio sample.

J/A+A/421/623/  table2.dat table3.dat  (24Kb)

Kaas A.A., Olofsson G., Bontemps S., Andre P., Nordh L., Huldtgren M., Prusti T., Persi P., Delgado A.J., Motte F., Abergel A., Boulanger F., Burgdorf M., Casali M.M., Cesarsky C.J., Davies J., Falgarone E., Montmerle T., Perault M., Puget J.L., Sibille F. Astron. Astrophys. 421, 623 (2004) The young stellar population in the Serpens Cloud Core: An ISOCAM survey.

J/A+A/421/643/  table1.dat  (14Kb)

Reyle C., Robin A.C. Astron. Astrophys. 421, 643 (2004) New nearby stars in the Liverpool-Edinburgh high proper motion survey selected by DENIS photometry.

J/A+A/421/763/  catalog.dat  (1.8 Mb)

Pokorny R.S., Jones H.R.A., Hambly N.C., Pinfield D.J. Astron. Astrophys. 421, 763 (2004) The Liverpool-Edinburgh High Proper Motion Catalogue.

J/A+A/421/913 See II/253/

J/A+A/421/937/  table7.dat  (24Kb)

Gratton, R.G., Bragaglia, A., Clementini, G., Carretta, E., Di Fabrizio, L., Maio, M., Taribello, E. Astron. Astrophys. 421, 937 (2004) Metal abundances of RR Lyrae stars in the bar of the Large Magellanic Cloud.

J/A+A/421/991/  table1.dat  (78Kb)

Piatti A.E., Claria J.J., Ahumada A.V. Astron. Astrophys., 421, 991-999 (2004) Improvements on the fundamental parameters of the open cluster Tombaugh 1 through Washington system photometry.

J/A+A/421/1087/  appena.dat appena.ps appena.tex table1.dat table1.ps table1.tex  (77Kb)

Harjunpaa P., Lehtinen K., Haikala L.K. Astron. Astrophys. 421, 1087 (2004) The relationship of CO abundance to extinction and N(H_{2}): Observations of globules and the dependence on star formation activity.

J/A+A/421/1121 See VI/113/

J/A+A/422/39/  table4.dat table6.dat  (412Kb)

Stickel M., Lemke D., Klaas U., Krause O., Egner S. Astron. Astrophys., 422, 39-54 (2004) The ISOPHOT 170μm Serendipity Survey. II. The catalog of optically identified galaxies.

J/A+A/422/129/  table1.dat  (45Kb)

Meynadier F., Heydari-Malayeri M., Deharveng L., Charmandaris V., Le Bertre T., Rosa M.R., Schaerer D., Zinnecker H. Astron. Astrophys. 422, 129 (2004) Stellar populations associated with the LMC Papillon Nebula.

J/A+A/422/141/  table2.dat  (17Kb)

Caratti o Garatti A., Giannini T., Lorenzetti D., Massi F., Nisini B., Vitali F. Astron. Astrophys. 422, 141 (2004) Physical and kinematical properties of a newly discovered star cluster and its associated jet in the Vela Molecular Clouds.

J/A+A/422/337/  table1.dat  (20Kb)

Tripathi D., Bothmer V., Cremades H. Astron. Astrophys. 422, 337 (2004) The Basic Characteristics of EUV Post-Eruptive Arcades and Their Role as Tracers of Coronal Mass Ejection Source Regions.

J/A+A/422/377/  table.dat  (47Kb)

Qiao R.C., Shen K.X., Harper D., Liu J.R. Astron. Astrophys., 422, 377-379 (2004) 1997-2000 CCD astrometric observations of Saturn’s satellites and comparison with theories.

J/A+A/422/407/  catalog.dat  (11.2 Mb)

Gavazzi R., Mellier Y., Fort B., Cuillandre J.-C., Dantel-Fort M. Astron. Astrophys. 422, 407 (2004) Mass and Light in the Supercluster of Galaxies MS0302+17.

J/A+A/422/511/  table3.dat  (9Kb)

Cedres B., Urbaneja M.A., Cepa J. Astron. Astrophys. 422, 511 (2004) The oxygen abundance gradient in M101: the reliability of the P method.

J/A+A/422/527/ table1.dat table2.dat table3.dat table4.dat table5.dat table6.dat table7.dat table8a.dat table8b.dat table9.dat (176Kb) Schuster W.J., Beers T.C., Michel R., Nissen P.E., Garcia G.

Astron. Astrophys. 422, 527 (2004) uvby-beta photometry of high-velocity and metal-poor stars. X. Stars of very low metal abundance: observations, reddenings, metallicities, classifications, distances, and relative ages.

J/A+A/422/627/  table1.dat table2.dat table3.dat tables.tex  (45Kb)

Balega I., Balega Y.Y., Maksimov A.F., Pluzhnik E.A., Schertl D., Shkhagosheva Z.U., Weigelt G. Astron. Astrophys. 422, 627 (2004) Speckle interferometry of nearby multiple stars. II.

J/A+A/422/731/  table1.dat table3.dat table4.dat  (147Kb)

Del Zanna G., Berrington K.A., Mason H.E. Astron. Astrophys., 422, 731-749 (2004) Benchmarking atomic data for astrophysics: Fe X.

J/A+A/422/907/  tab2-4.dat table5.dat  (17Kb)

Jaeger K., Ziegler B.L., Boehm A., Heidt J., Moellenhoff C., Hopp U., Mendez R.H., Wagner S. Astron. Astrophys., 422, 907-913 (2004) Internal kinematics of spiral galaxies in distant clusters. II. Observations and data analysis.

J/A+A/422/925/  appenda.txt table2.dat table5.dat  (29Kb)

Shapovalova A.I., Doroshenko V.T., Bochkarev N.G., Burenkov A.N., Carrasco L., Chavushyan V.H., Collin S., Valdes J.R., Borisov N., Dumont A.-M., Vlasuyk V.V., Chilingarian I., Fioktistova I.S., Martinez O.M. Astron. Astrophys. 422, 925 (2004) Profile variability of the Hαand Hβbroad emission lines in NGC 5548

J/A+A/422/941/  refs.dat table1.dat  (114Kb)

Casasola V., Bettoni D., Galletta G. Astron. Astrophys. 422, 941 (2004) The gas content of peculiar galaxies: strongly interacting systems

J/A+A/422/1023/  table4.dat table5.dat  (12Kb)

Strigachev A., Lampens P. Astron. Astrophys. 422, 1023 (2004) Multicolour CCD measurements of nearby visual double stars. II.

J/A+A/422/1109/  tablea1.dat tablea2.dat  (13Kb)

Colon C., Alonso Medina A. Astron. Astrophys. 422, 1109 (2004) Lifetimes and oscillator strengths for the 5s5p6s,5s5p5d and 5p3 levels in single ionized tin.

J/A+A/423/97/  macho.dat ogle.dat  (11Kb)

Borissova J., Minniti D., Rejkuba M., Alves D., Cook H.K., Freeman K.C. Astron. Astrophys. 423, 97 (2004) Properties of RR Lyrae stars in the Inner Regions of the Large Magellanic Cloud.

J/A+A/423/133/  position.dat tablea1.dat  (27Kb)

Westera P., Cuisinier F., Telles E., Kehrig C. Astron. Astrophys., 423, 133-146 (2004) Stellar populations in HII galaxies.

J/A+A/423/189/  table1.dat table2.dat  (47Kb)

Eigenbrod A., Mermilliod J.-C., Claria J.J., Andersen J., Mayor M. Astron. Astrophys. 423, 189 (2004) Red giants in open clusters: XI. Membership, duplicity and structure of NGC 2477

J/A+A/423/199/  table1.dat table2.dat  (24Kb)

Costa R.D.D., Uchida M.M.M., Maciel W.J. Astron. Astrophys. 423, 199 (2004) Chemical abundances of planetary nebulae towards the galactic anticenter.

J/A+A/423/235/  table3.dat  (16Kb)

Chandra S., Shine S.A. Astron. Astrophys. 423, 235 (2004) Suggestions for an interstellar C_{5}H_{2} search

J/A+A/423/353/  table6.dat table6.ps table6.tex  (52Kb)

Jasniewicz G., de Laverny P., Parthasarathy M., Lebre A., Thevenin F. Astron. Astrophys. 423, 353 (2004) On the nature of UV-bright stars in globular clusters from VLT-UVES spectra.

J/A+A/423/381/  table2.dat table3.dat table4.dat  (138Kb)

Baudrand A., Bec-Borsenberger A., Borsenberger J.

Astron. Astrophys. 423, 381 (2004) Asteroidal I, J, K magnitudes recovered in the DENIS Survey: second release.

J/A+A/423/397/  table1.dat table2.dat  (29Kb)

Ben Nessib N., Dimitrijevic M.S., Sahal-Brechot S. Astron. Astrophys. 423, 397 (2004) Stark broadening of the four times ionized silicon spectral lines.

J/A+A/423/449/  table1.dat table2.dat table3.dat  (35Kb)

Popesso P., Bohringer H., Brinkmann J., Voges W., York D.G. Astron. Astrophys. 423, 449 (2004) RASS-SDSS Galaxy Clusters Survey. I. The Catalog and the correlation of X-ray and optical properties.

J/A+A/423/469/  table1.dat table2.dat  (21Kb)

Sazonov S.Yu., Revnivtsev M.G. Astron. Astrophys. 423, 469 (2004) Statistical properties of local active galactic nuclei inferred from the RXTE 3-20 keV All-Sky Survey.

J/A+A/423/517/  table1.dat  (39Kb)

Rocha-Pinto H.J., Flynn C., Hanninen J., Scalo J., Maciel W.J., Hensler G. Astron. Astrophys. 423, 517 (2004) Chemical enrichment and star formation in the Milky Way disk. III. Chemodynamical constraints.

J/A+A/423/579/  table2.dat  (16Kb)

Mizutani M., Onaka T., Shibai H. Astron. Astrophys., 423, 579-592 (2004) Origin of diffuse [CII]158 micron and [SiII]35 micron emission in the Carina nebula.

J/A+A/423/677/  refs.dat table2.dat  (20Kb)

Morel T., Micela G. Astron. Astrophys. 423, 677 (2004) On the determination of oxygen abundances in chromospherically active stars.

J/A+A/423/683/  table2.dat  (9Kb)

Shi J.R., Gehren T., Zhao G. Astron. Astrophys., 423, 683-691 (2004) Sodium abundances in nearby disk stars.

J/A+A/423/755/  table1.dat  (78Kb)

Pourbaix D., Ivezic Z., Knapp G.R., Gunn J.E., Lupton R.H. Astron. Astrophys. 423, 755 (2004) Color-Induced Displacement double stars in SDSS.

J/A+A/423/849/  table1.dat table2.dat  (11Kb)

Grosbol P., Patsis P.A., Pompei E. Astron. Astrophys. 423, 849 (2004) Spiral galaxies observed in the near-infrared K band. I. Data analysis and structural parameters.

J/A+A/423/867/  table4.dat table5.dat table6.dat table7.dat table12.dat  (41Kb)

Liang Y.C., Hammer F., Flores H., Elbaz D., Marcillac D., Cesarsky C.J. Astron. Astrophys., 423, 867-880 (2004) The Luminosity-Metallicity relation of distant luminous infrared galaxies.

J/A+A/423/919/  table2.dat table3.dat  (23Kb)

Bratsolis E., Kontizas M., Bellas-Velidis I. Astron. Astrophys., 423, 919-924 (2004) Triggered star formation in the inner wing of the SMC. Two possible supernova explosions in the N83-84-85 region.

J/A+A/423/935/  tablea1.dat  (44Kb)

Massaro E., Mantovani F., Fanti R., Maesano M., Montagni F., Nesci R. Sclavi S., Tosti G., Venturi T. Astron. Astrophys. 423, 935 (2004) VLBI imaging and optical variability of the BL Lac object OQ 530 (B1418+546).

J/A+A/423/983/  table3.dat  (6Kb)

Mutschke H., Andersen A.C., Jaeger A.C., Henning T., Braatz A. Astron. Astrophys. 423, 983 (2004) Optical data of meteoritic nano-diamonds from far-ultraviolet to far-infrared wavelengths.

J/A+A/423/1029/  refs.dat table2.dat table4.dat table35.dat  (30Kb)

Stelzer B., Micela G., Neuhaeuser R. Astron. Astrophys., 423, 1029-1044 (2004)

XMM-Newton probes the stellar population in Chamaeleon I South.

J/A+A/423/1109/  sp/* table1.dat  (1.0 Mb)

Allende Prieto C., Asplund M., Fabiani Bendicho P. Astron. Astrophys., 423, 1109-1117 (2004) Center-to-limb variation of solar line profiles as a test of NLTE line formation calculations.

J/A+A/424/L31/  table2.dat  (13Kb)

Moutou C., Pont F., Bouchy F., Mayor M. Astron. Astrophys. 424, L31 (2004) Accurate radius and mass of the transiting exoplanet OGLE-TR-132b.

J/A+A/424/79/  table1.dat table4.dat  (36Kb)

Mercurio A., Busarello G., Merluzzi P., La Barbera F., Girardi M., Haines C. Astron. Astrophys. 424, 79 (2004) Transformations of galaxies in the environments of the cluster ABCG 209 at z 0.2

J/A+A/424/125/  table2.dat  (60Kb)

Demers S., Battinelli P., Letarte B Astron. Astrophys. 424, 125 (2004) A carbon star approach to IC 10: Distance and correct size.

J/A+A/424/272 See V/122/

J/A+A/424/363/  table6.dat table7.dat table8.dat  (5.2 Mb)

Jonauskas V., Keenan F.P., Foord M.E., Heeter R.F., Rose S.J., Ferland G.J., Kisielius R., van Hoof P.A.M., Norrington P.H. Astron. Astrophys. 424, 363 (2004) Dirac-Fock energy levels and transition probabilities for oxygen-like Fe XIX.

J/A+A/424/371/  table1.dat  (66Kb)

Morganti R. Garrett M.A., Chapman S., Baan W., Helou G., Soifer T. Astron. Astrophys. 424, 371 (2004) A Deep WSRT 1.4 GHz Radio Survey of the Spitzer Space Telescope FLSv Region

J/A+A/424/447/  galaxies.dat table3.dat table4.dat  (118Kb)

Pizzella A., Corsini E.M., Vega-Beltran J.C., Bertola F. Astron. Astrophys. 424, 447 (2004) Ionized gas and stellar kinematics of seventeen nearby spiral galaxies.

J/A+A/424/519/  table1.dat  (10Kb)

Hawkins M.R.S. Astron. Astrophys., 424, 519-529 (2004) Naked active galactic nuclei.

J/A+A/424/545/  table1.dat table2.dat table4.dat  (13Kb)

Padovani P., Allen M.G., Rosati P., Walton N.A. Astron. Astrophys., 424, 545-559 (2004) Discovery of optically faint obscured quasars with Virtual Observatory tools.

J/A+A/424/727 See V/122/

J/A+A/424/919/  grid.dat tables.dat x70z02/*  (34.7 Mb)

Claret A. Astron. Astrophys. 424, 919 (2004) New grids of stellar models including tidal-evolution constants up to carbon burning. I. From 0.8 to 125M_{⊙} at Z=0.02.

J/A+A/424/993/ (2Kb) Kjurkchieva D.P., Marchev D.V., Heckert P.A., Shower C.A. Astron. Astrophys., 424, 993-1002 (2004) Spectroscopic and photometric observations of the short-period RS CVn-type star BH Virginis.

J/A+A/424/1101/  blc/* fig2/* table1.dat vlc/*  (7.4 Mb)

Kaluzny J., Olech A., Thompson I.B., Pych W., Krzeminski W., Szchwarzenberg-Czerny A. Astron. Astrophys. 424, 1101 (2004) Cluster AgeS Experiment (CASE). Catalog of variable stars in the field of globular cluster Omega Centauri.

J/A+A/425/99/  table5.dat table6.dat  (11Kb)

Lal D.V., Shastri P., Gabuzda D.C. Astron. Astrophys. 425, 99 (2004) Milliarcsec-scale radio structure of a matched sample of Seyfert 1 and Seyfert 2 galaxies.

J/A+A/425/187/  table1.dat table2.dat table6.dat  (28Kb)

Thoren P., Edvardsson B., Gustafsson B. Astron. Astrophys. 425, 187 (2004) Subgiants as probes of galactic chemical evolution.

J/A+A/425/263/  gf/* linelist.txt obsspect.dat synspect.dat  (52.8 Mb)

Castelli F., Hubrig S. Astron. Astrophys., 425, 263-270 (2004) A spectroscopic atlas of the HgMn star HD 175640 (B9 V) λλ3040-10000Å.

J/A+A/425/367/ apx.txt reflex.dat reflex50.dat reflex70.dat refs.dat table2.dat table3.dat table4.dat table5.dat table8.dat table9.dat table10.dat table11.dat table12.dat (1.0 Mb) Boehringer H., Schuecker P., Guzzo L., Collins C.A., Voges W., Cruddace R.G., Ortiz-Gil A., Chincarini G., De Grandi S., Edge A.C., MacGillivray H.T., Neumann D.M., Schindler S., Shaver P. Astron. Astrophys. 425, 367 (2004) The ROSAT-ESO Flux Limited X-ray (REFLEX) Galaxy Cluster Survey. V. The cluster catalogue.

J/A+A/425/417/  table1.dat  (15Kb)

Kong Xu Astron. Astrophys., 425, 417-427 (2004) Spectroscopic study of blue compact galaxies. IV. Star formation rates and gas depletion timescales.

J/A+A/425/429/  table7.dat  (11Kb)

Cortese L., Gavazzi G., Boselli A., Iglesias-Paramo J., Carrasco L. Astron. Astrophys. 425, 429 (2004) Multiple merging in the Abell cluster 1367.

J/A+A/425/489/  sp/* table1.dat  (953Kb)

Comeron F., Torra J., Chiappini C., Figueras F., Ivanov V.D., Ribas S.J. Astron. Astrophys. 425, 489 (2004) A search for late-type supergiants in the inner regions of the Milky Way.

J/A+A/425/509/  catalog.dat  (1.1 Mb)

Andreuzzi G., Testa V., Marconi G., Alcaino G., Alvarado F., Buonanno R. Astron. Astrophys., 425, 509-518 (2004) VLT FORS-1 observations of NGC 6397: Evidence for mass segregation.

J/A+A/425/595/ refs.dat table2.dat table2.tex table3.dat table3.tex table6.dat table6.tex table7.dat table7.tex table8.dat table8.tex (1.9 Mb) Groenewegen M.A.T. Astron. Astrophys. 425, 595 (2004) Long Period Variables in the Magellanic Clouds: OGLE + 2MASS + DENIS.

J/A+A/425/671/  ewline.dat stars.dat  (29Kb)

Bai G.S., Zhao G., Chen Y.Q., Shi J.R., Klochkova V.G., Panchuk V.E., Qiu H.M., Zhang H.W. Astron. Astrophys. 425, 671 (2004) Chemical abundances of 10 metal-poor halo stars.

J/A+A/425/767/ table5.dat table6.dat table7.dat table8.dat table9.dat table10.dat table11.dat (30Kb) Yamada K.M.T., Degli Esposti C., Botschwina P., Foerster P., Bizzocchi L., Thorwirth S., Winnewisser G. Astron. Astrophys. 425, 767 (2004) Millimeter-wave spectroscopy of HC_{5}N in vibrationally excited states below 500cm^{−1}.

J/A+A/425/797/  table2.dat  (38Kb)

La Barbera F., Merluzzi P., Busarello G., Massarotti M., Mercurio A. Astron. Astrophys. 425, 797 (2004) Probing galaxy evolution through the internal colour gradients, the Kormendy relations and the Photometric Plane of cluster galaxies at z 0.2.

J/A+A/425/927/  table1.dat  (558Kb)

Klotz A., Caux E., Monin J.-.L, Lodieu N. Astron. Astrophys. 425, 927 (2004) Catalog of objects in the Serpens cloud field.

J/A+A/425/1107/  catalog.dat  (10Kb)

Vieira Martins R., Veiga C. H., Bourget P., Andrei A. H., Descamps P. Astron. Astrophys. 425, 1107 (2004) Systematic astrometric observations of Proteus.

J/A+A/425/1147/  table1.dat  (16Kb)

Dimitrijevic M.S., Dacic M., Cvetkovic Z., Simic Z.

Astron. Astrophys. 425, 1147 (2004) Stark broadening of Ga I spectral lines.

J/A+A/426/11/  table3.dat  (232Kb)

Pietsch W., Misanovic Z., Haberl F., Hatzidimitriou D., Ehle M., Trinchieri G. Astron. Astrophys. 426, 11 (2004) XMM-Newton survey of the Local Group galaxy M 33.

J/A+A/426/81/  fig7a.fit fig7aerr.fit fig7b.fit fig7berr.fit  (623Kb)

Paumard T., Maillard J.-P., Morris M. Astron. Astrophys. 426, 81 (2004) Kinematic and structural analysis of the Minispiral in the Galactic Center from BEAR spectro-imagery.

J/A+A/426/97/  table1.dat  (16Kb)

Faundez S., Bronfman L., Garay G., Chini R., Nyman L.-A., May J. Astron. Astrophys., 426, 97-103 (2004) SIMBA survey of southern high-mass star forming regions. I. Physical parameters of the 1.2-mm/IRAS sources.

J/A+A/426/119/  table1.dat table2.dat  (16Kb)

Mookerjea B., Kramer C., Nielbock M., Nyman L.-A. Astron. Astrophys., 426, 119-129 (2004) The Giant Molecular Cloud associated with RCW 106. A 1.2 mm continuum mapping study.

J/A+A/426/171/  table2.dat  (13Kb)

Khanzadyan T., Gredel R., Smith M.D., Stanke T. Astron. Astrophys. 426, 171 (2004) An unbiased search for the signatures of protostars in the rho Ophiuchi A Molecular Cloud. I. Near-infrared observations.

J/A+A/426/247/  table1.dat table2.dat  (14Kb)

Chapellier E., Mathias P., Garrido R., Le Contel J.-M., Sareyan J.-P., Ribas I., Parrao L., Moya A., Pena J.H., Alvarez M. Astron. Astrophys. 426, 247 (2004) HD 173977: an ellipsoidal delta Scuti star variable.

J/A+A/426/367/  table1.dat  (93Kb)

Mickaelian A.M. Astron. Astrophys. 426, 367 (2004) DSS1/DSS2 astrometry for 1101 First Byurakan Survey blue stellar objects: accurate positions and other results.

J/A+A/426/463/  table1.dat table2.dat tables.ps tables.tex  (56Kb)

Orienti M., Dallacasa D., Fanti C., Fanti R., Tinti S., Stanghellini C. Astron. Astrophys. 426, 463 (2004) The B3-VLA CSS sample. V. VLBA images at 6 and 3.6 cm.

J/A+A/426/503/  notes.dat refs.dat table1a.dat table1b.dat  (228Kb)

Wu Y., Wei Y., Zhao M., Shi Y., Yu W., Qin S., Huang M. Astron. Astrophys. 426, 503 (2004) A study of high velocity molecular outflows with an up-to-date sample.

J/A+A/426/695/  table2.dat  (8Kb)

Zucker S., Mazeh T., Santos N.C., Udry S., Mayor M. Astron. Astrophys. 426, 695 (2004) Multi-order TODCOR: application to observations taken with the CORALIE echelle spectrograph. II. A planet in the system HD41004.

J/A+A/426/717/  table1.dat table2.dat table3.dat table4.dat tables.tex  (49Kb)

Tayal S.S. Astron. Astrophys. 426, 717 (2004) Accurate calculation of oscillator strengths for Cl II lines using non-orthogonal wavefunctions.

J/A+A/426/819/  table5.dat  (68Kb)

Balaguer-Nunez L., Jordi C., Galadi-Enriquez D., Zhao J.L. Astron. Astrophys. 426, 819 (2004) New membership determination and proper motions of NGC 1817. Parametric and non-parametric approach.

J/A+A/426/827/  table4.dat table5.dat  (1.6 Mb)

Balaguer-Nunez L., Jordi C., Galadi-Enriquez D., Masana E. Astron. Astrophys. 426, 827 (2004) uvby-HβCCD Photometry of NGC 1817 and NGC 1807.

J/A+A/426/1001/  table2a.dat table2b.dat table2c.dat table2d.dat table3.dat  (87Kb)

Csizmadia Sz., Klagyivik P. Astron. Astrophys. 426, 1001 (2004) On the properties of contact binary stars.

J/A+A/426/1007/  tablea1.dat  (19Kb)

Morel T., Micela G., Favata F., Katz D., Pillitteri I. Astron. Astrophys. 426, 1007 (2004) The photospheric abundances of active binaries. III. Abundance peculiarities at high activity level.

J/A+A/426/1135 See VI/112/

J/A+A/427/23/  lhss cov.fit lhss map.fit lhss snr.fit table3.dat table4.dat  (22.9 Mb)

Fadda D., Lari C., Rodighiero G., Franceschini A., Elbaz D., Cesarsky C. Perez-Fournon I. Astron. Astrophys. 427, 23 (2004) ISOCAM observations in the Lockman Hole. I. The 14.3 micron shallow survey: data reduction, catalogue, and optical identifications.

J/A+A/427/87/  list.dat  (25Kb)

Gorosabel J., Lund N., Brandt S., Westergaard N.J., Castro Ceron J.M. Astron. Astrophys., 427, 87-93 (2004) The potential of INTEGRAL for the detection of high redshift GRBs.

J/A+A/427/107/  table1.dat  (31Kb)

Lamy H., Hutsemekers D. Astron. Astrophys. 427, 107 (2004) Polarization properties of broad absorption line QSOs : new statistical clues.

J/A+A/427/231/  table1.dat table2.dat table3.dat table3.txt  (185Kb)

Gorny S.K., Stasinska G., Escudero A.V., Costa R.D.D. Astron. Astrophys., 427, 231-244 (2004) The populations of planetary nebulae in the direction of the Galactic Bulge. Chemical abundances and Wolf-Rayet central stars.

J/A+A/427/313/  table1.dat table2.dat table3.dat table4.dat  (79Kb)

De Medeiros J.R., Udry S., Mayor M. Astron. Astrophys., 427, 313-317 (2004) A catalog of rotational and radial velocities for evolved stars. III. Double-lined binary systems.

J/A+A/427/371/  table4.dat table5.dat table6.dat table7.dat tables.ps tables.tex  (76Kb)

Lainey V., Arlot J.E., Vienne A Astron. Astrophys. 427, 371 (2004) New accurate ephemerides for the Galilean satellites of Jupiter. II. Fitting the observations.

J/A+A/427/387/  freelunc.dat knownobj.dat master.dat qsocand.dat  (322.2 Mb)

Flesch E., Hardcastle M.J. Astron. Astrophys. 427, 387 (2004) An all-sky optical catalogue of radio / X-ray sources.

J/A+A/427/397/  cl1040.dat cl1216.dat cl1232.dat cl105411.dat cl105412.dat  (57Kb)

Halliday C., Milvang-Jensen B., Poirier S., Poggianti B.M., Jablonka P., Aragon-Salamanca A., Saglia R.P., De Lucia G., Pello R., Simard L., Clowe D.I., Rudnick G., Dalcanton J.J., White S.D.M., Zaritsky D. Astron. Astrophys. 427, 397 (2004) Spectroscopy of clusters in the ESO Distant Cluster Survey (EDisCS): Redshifts, velocity dispersions and substructure for 5 clusters.

J/A+A/427/613/  tablea1.dat  (146Kb)

Kerschbaum F., Nowotny W., Olofsson H., Schwarz H.E. Astron. Astrophys., 427, 613-619 (2004) A census of AGB stars in Local Group galaxies. III. The dwarf spheroidal And II.

J/A+A/427/763/  table1.dat table2.dat table3.dat tables.ps tables.tex  (850Kb)

Aggarwal K.M., Keenan F.P. Astron. Astrophys. 427, 763 (2004) Radiative rates for E1, E2, M1 and M2 transitions in Fe X.

J/A+A/427/769/  table2.dat table31.dat  (850Kb)

Terasranta H., Achren J., Hanski M., Heikkila J., Holopainen J., Joutsamo O., Juhola M., Karlamaa K., Katajainen S., Keinanen P., Koivisto P., Koskimies M., Kononen P., Lainela M., Lahtenmaki A., Makinen K., Niemela T., Nurmi P.,T.Pursimo, Rekola R., Savolainen T., Tornikoski M., Torppa J., Valtonen T., Varjonen N., Vilenius E., Virtanen J., Wiren S.

Astron. Astrophys. 427, 769 (2004) Twenty years monitoring of extragalactic sources at 22, 37 and 87 GHz.

J/A+A/427/773/  table2.dat  (34Kb)

Rodighiero G., Lari C., Fadda D., Franceschini A., Elbaz D., Cesarsky C. Astron. Astrophys. 427, 773 (2004) ISOCAM observations in the Lockman Hole. II. The 14.3μm deep survey: data reduction, catalogue and source counts.

J/A+A/427/933/  table1.dat  (17Kb)

Kovtyukh V.V., Soubiran C., Belik S.I. Astron. Astrophys., 427, 933-936 (2004) A new Boehm-Vitense gap in the temperature range 5560 to 5610K in the main sequence.

J/A+A/428/67/  table1.dat table2.dat table3.dat  (23Kb)

Villanova S., Carraro G., de la Fuente Marcos R., Stagni R. Astron. Astrophys. 428, 67 (2004) NGC 5385, NGC 2664 and Collinder 21: three candidate Open Cluster Remnants.

J/A+A/428/149/  refs.dat table1.dat table2.dat  (17Kb)

Schroeder S.E., Kaper L., Lamers H.J.G.L.M., Brown A.G.A. Astron. Astrophys., 428, 149-157 (2004) On the Hipparcos parallaxes of O stars.

J/A+A/428/339/  tablea1.dat tablea2.dat tablea3.dat  (54Kb)

Alcala J.M., Pannella M., Puddu E., Radovich M., Silvotti R., Arnaboldi M., Capaccioli M., Covone G., Dall’Ora M., De Lucia G., Grado A., Longo G., Mercurio A., Musella I., Napolitano N., Pavlov M., Rifatto A., Ripepi V., Scaramella R. Astron. Astrophys. 428, 339 (2004) The Capodimonte Deep Field: Presentation of the survey and first follow-up studies.

J/A+A/428/383/  table1.dat table2.dat table3.dat table4.dat tableb1.dat  (84Kb)

Della Ceca R., Maccacaro T., Caccianiga A., Severgnini P., Braito V., Barcons X., Carrera F.J., Watson M.G., Tedds J.A., Brunner H., Lehmann I., Page M.J., Lamer G., Schwope A. Astron. Astrophys., 428, 383-399 (2004) Exploring the X-ray sky with the XMM-Newton Bright Serendipitous Survey.

J/A+A/428/555/  table1.dat table2.dat table4.dat table5.dat tables.ps tables.tex  (49Kb)

Sollerman J., Lindahl J., Kozma C., Challis P., Filippenko A.V., Fransson C., Garnavich P.M., Leibundgut B., Li W., Lundqvist P., Milne P., Spyromilio J., Kirshner R.P. Astron. Astrophys. 428, 555 (2004) The late-time light curve of the Type Ia supernova 2000cx.

J/A+A/428/587/  stars.dat table3.dat  (23Kb)

Kervella, P., Bersier, D., Mourard, D., Nardetto, N., Fouque, P., Coude du Foresto, V. Astron. Astrophys. 428, 587 (2004) Cepheid distances from infrared long-baseline interferometry. III. Calibration of the surface brightness-color relations.

J/A+A/428/823/  appen.dat galaxies.dat table1.dat table2.dat table3.dat  (44Kb)

O’Neil K.O., Bothun G., van Driel W., Monnier-Ragnine D. Astron. Astrophys. 428, 823 (2004) A new HI catalog of low surface brightness galaxies out to z=0.1 - Tripling the number of massive LSB galaxies known.

J/A+A/428/837/  galaxies.dat table4.dat table5.dat table6.dat tables.ps tables.tex  (103Kb)

Ferreiro D.L., Pastoriza M.G. Astron. Astrophys. 428, 837 (2004) Sample of minor mergers of galaxies: Optical CCD surface photometry and HII region properties.

J/A+A/428/877/  tablea1.dat tablea2.dat  (12Kb)

Krajnovic D., Jaffe W. Astron. Astrophys. 428, 877 (2004) HST observations of nuclear stellar disks.

J/A+A/428/1001/ table2.dat table3.dat table4.dat table5.dat table6.dat table7.dat table8.dat table9.dat table10.dat table11.dat (7.6 Mb) Claret A. Astron. Astrophys. 428, 1001 (2004) A new non-linear limb-darkening law for LTE stellar atmosphere models. III. Sloan filters: Calculations for -5.0 =log[M/H] =+1, 2000K =Teff =50000K at several surface gravities.

J/A+A/428/1019/  refs.dat table1.dat table2.dat table3.dat tables.tex  (42Kb)

Mueller H.S.P., Menten K.M., Maeder H. Astron. Astrophys., 428, 1019-1026 (2004) Accurate rest frequencies of methanol maser and dark cloud lines.

J/A+A/428/1043/  vcdfs.dat  (111Kb)

Le Fevre O., Vettolani G., Paltani S., Tresse L., Zamorani G., Le Brun V., Moreau C., Bottini D., Maccagni D., Picat J.P., Scaramella R., Scodeggio M., Zanichelli A., Adami C., Arnouts S., Bardelli S., Bolzonella M., Cappi A., Charlot S., Contini T., Foucaud S., Franzetti P., Garilli B., Gavignaud I., Guzzo L., Ilbert O., Iovino A., Mccracken H.J., Mancini D., Marano B., Marinoni C., Mathez G., Mazure A., Meneux B., Merighi R., Pello R., Pollo A., Pozzetti L., Radovich M., Zucca E., Arnaboldi M., Bondi M., Bongiorno A., Busarello G., Ciliegi P., Gregorini L., Mellier Y., Merluzzi P., Ripepi V., Rizzo D. Astron. Astrophys., 428, 1043-1049 (2004) The VIMOS VLT Deep Survey. Public release of 1599 redshifts to I_{AB} =24 across the Chandra Deep Field South.

J/A+A/429/101/  table2.dat table4.dat table8.dat table56.dat  (25Kb)

Serote Roos M., Lobo C., Durret F., Iovino A., Marquez I. Astron. Astrophys., 429, 101-114 (2005) The intermediate-redshift galaxy cluster CL 0048-2942. Stellar populations.

J/A+A/429/115/  table1.dat table2.dat  (16Kb)

Noeske K.G., Papaderos P., Cairos L.M., Fricke K.J. Astron. Astrophys. 429, 115 (2005) New insights to the photometric structure of Blue Compact Dwarf Galaxies from deep near-infrared studies. II. The sample of northern BCDs.

J/A+A/429/235/  refs.dat table3.dat  (13Kb)

Bergeat J., Chevallier L. Astron. Astrophys. 429, 235 (2005) The mass loss of C-rich giants.

J/A+A/429/309/  table1.dat table2.dat tables.ps tables.tex  (88Kb)

Frasca A., Lanza A.F. Astron. Astrophys. 429, 309 (2005) Orbital period variation in close binaries from radial velocity data and magnetic activity cycles. II. HR 1099.

J/A+A/429/317/  table2.dat table3.dat table4.dat table6.dat tablea.dat  (26Kb)

Adelman S.J., Young K.J. Astron. Astrophys. 429, 317 (2005) uvby FCATP Photometry of the HgMn star 46 (rho) Aqr and the mCP stars HR 2258, MW Vul, and HR 9017.

J/A+A/429/559/  fig2.ps fig4.ps table3.dat  (191Kb)

Sollerman J., Cox N., Mattila S., Ehrenfreund P., Kaper L., Leibundgut B., Lundqvist P. Astron. Astrophys. 429, 559 (2005) Diffuse Interstellar Bands in NGC 1448.

J/A+A/429/607/  table2a.dat table2b.dat table2c.dat  (162Kb)

Ortolani S., Bica E., Barbuy B., Zoccali M. Astron. Astrophys. 429, 607 (2005) The old open clusters Berkeley 36, Berkeley 73 and Biurakan 13 (Berkeley 34).

J/A+A/429/819/  residual.dat table4.dat  (31Kb)

Michard R. Astron. Astrophys., 429, 819-823 (2005) Peculiarities and populations in elliptical galaxies. II. Visual-near IR colours as population indices.

J/A+A/429/837/  table1.dat table2.dat  (1.6 Mb)

Cioni M.-R.L., Habing H.J. Astron. Astrophys. 429, 837 (2005) Near-IR observations of NGC 6822: AGB stars, distance, metallicity and structure.

J/A+A/429/963/  table1.dat table2.dat  (27Kb)

Ozawa H., Grosso N., Montmerle T. Astron. Astrophys. 429, 963 (2005) The X-ray emission from Young Stellar Objects in the rho Ophiuchi cloud core as seen by XMM-Newton.

J/A+A/429/1007/  table2.dat  (15Kb)

Scholz A., Eisloffel J. Astron. Astrophys. 429, 1007 (2005)

Rotation and variability of very low mass stars and brown dwarfs near epsilon Ori.

J/A+A/429/1117/  table1.dat table2.dat table3.dat tables.ps tables.tex  (3.8 Mb)

Aggarwal K.M., Keenan F.P. Astron. Astrophys. 429, 1117 (2005) Effective collision strengths for transitions in Fe XIII

J/A+A/429/1125/  notes.dat table1.ps table1a.dat table1b.dat  (252Kb)

Liu Q.Z., Mirabel I.R. Astron. Astrophys. 429, 1125 (2005) A catalogue of ultraluminous X-ray sources in external galaxies.

J/A+A/430/19/  tablea.dat  (27Kb)

Ferrari C., Benoist C., Maurogordato S., Cappi A., Slezak E. Astron. Astrophys. 430, 19 (2005) Dynamical state and star formation properties of the merging galaxy cluster Abell 3921.

J/A+A/430/115/  table1.dat  (21Kb)

Hammer F., Flores H., Elbaz D., Zheng X.Z., Liang Y.C., Cesarsky C. Astron. Astrophys., 430, 115-128 (2005) Did most present-day spirals form during the last 8 Gyr? A formation history with violent episodes revealed by panchromatic observations.

J/A+A/430/137/  table1.dat  (28Kb)

Kouwenhoven M.B.N., Brown A.G.A., Zinnecker H., Kaper L., Portegies Zwart S.F. Astron. Astrophys., 430, 137-154 (2005) The primordial binary population. I. A near-infrared adaptive optics search for close visual companions to A star members of Scorpius OB2.

J/A+A/430/165/  notes.dat tablea1.dat  (1.6 Mb)

Famaey B., Jorissen A., Luri X., Mayor M., Udry S., Dejonghe H., Turon C. Astron. Astrophys. 430, 165 (2005) Local Kinematics of K and M Giants from CORAVEL/Hipparcos/Tycho-2 Data. Revisiting the concept of superclusters.

J/A+A/430/223/  table1.dat  (14Kb)

Lisker T., Heber U., Napiwotzki R., Christlieb N., Han Z., Homeier D., Reimers D. Astron. Astrophys., 430, 223-243 (2005) Hot subdwarfs from the ESO Supernova Ia Progenitor Survey. I. Atmospheric parameters and cool companions of sdB stars.

J/A+A/430/255/  table1.dat table2.dat  (74Kb)

Yushchenko A., Gopka V., Goriely S., Musaev F., Shavrina A., Kim C., Woon Kang Y., Kuznietsova J., Yushchenko V. Astron. Astrophys., 430, 255-262 (2005) Thorium-rich halo star HD 221170: Further evidence against the universality of the r-process.

J/A+A/430/287/  table1.dat table2.dat  (20Kb)

Marino A., Micela G., Peres G., Pillitteri I., Sciortino S. Astron. Astrophys., 430, 287-301 (2005) X-ray spectral and timing characteristics of the stars in the young open cluster IC 2391.

J/A+A/430/303/  table1.dat table2.dat table3.dat  (16Kb)

Cardini D. Astron. Astrophys., 430, 303-311 (2005) Mg II chromospheric radiative loss rates in cool active and quiet stars.

J/A+A/430/327/  cpasj3.dat  (726Kb)

Lu L.Z., Manrique W.T., Perdomo R., Wang Z.Z., Podesta R.C., Actis E.L., Zeng F.M., Alonso E., Zhao G., Pacheco A.M., Qiao Q.Y., Wang H.Q. Astron. Astrophys. 430, 327 (2005) Third San Juan photoelectric astrolabe catalogue (CPASJ3).

J/A+A/430/331/ (2Kb) Chidichimo M.C., Del Zanna G., Mason H.E., Badnell N.R., Tully J.A., Berrington K.A. Astron. Astrophys., 430, 331-341 (2005) Atomic data from the IRON Project. LVI. Electron excitation of Be-like Fe XXIII for the n=2,3,4 configurations.

J/A+A/430/373/  sp/* table2.dat table3.dat  (10.0 Mb)

Theureau G., Coudreau N., Hallet N., Hanski M., Alsac L., Bottinelli L., Gouguenheim L., Martin J.-M., Paturel G. Astron. Astrophys. 430, 373 (2005) 21-cm line measurements of 586 galaxies with the new Nancay receiver.

J/A+A/430/411/  table2.dat  (21Kb)

Gavazzi G., Donati A., Cucciati O., Sabatini S., Boselli A., Davies J., Zibetti S. Astron. Astrophys. 430, 411 (2005) The structure of elliptical galaxies in the Virgo cluster. Results from the INT Wide Field Survey.

J/A+A/430/421/  table1.dat  (48Kb)

Subramaniam, A. Astron. Astrophys. 430, 421 (2005) Reddening map of the Large Magellanic Cloud bar region.

J/A+A/430/471/  table2.dat  (202Kb)

Vazquez R.A., Baume G.L., Feinstein C., Nunez J.A., Vergne M.M. Astron. Astrophys., 430, 471-480 (2005) Revisiting the area of the open cluster Stock 16.

J/A+A/430/549/  sp/* spectra.dat table2.dat table3.dat  (1.7 Mb)

Wouterloot J.G.A., Brand J., Henkel C. Astron. Astrophys., 430, 549-560 (2005) The interstellar C^{18}O/C^{17}O ratio in the solar neighbourhood: The ρOphiuchus cloud.

J/A+A/430/603/ lc/* notes.dat table5.dat table6.dat table9.dat table10.dat table11.dat table14.dat (778Kb) Di Fabrizio L., Clementini G., Maio M., Bragaglia A., Carretta E., Gratton R., Montegriffo P., Zoccali M. Astron. Astrophys. 430, 603 (2005) Variable stars in the bar of the Large Magellanic Cloud: the photometric catalogue.

J/A+A/430/725/  table1.dat table2.dat  (98Kb)

Hudson C.E., Bell K.L. Astron. Astrophys. 430, 725 (2005) Effective collision strengths for fine-structure transitions for the electron impact excitation of N II.

J/A+A/430/751 See VII/242/

J/A+A/430/911/  notes.dat tables.dat  (15Kb)

Relano M., Beckman J.E. Astron. Astrophys., 430, 911-926 (2005) Expansive components in HII regions.

J/A+A/430/941/  table4.dat table5.dat  (202Kb)

Prisinzano L., Damiani F., Micela G., Sciortino S. Astron. Astrophys., 430, 941-957 (2005) The star formation region NGC 6530: Distance, ages and initial mass function.

J/A+A/430/1005/  table1.dat  (50Kb)

Lamm M.H., Mundt R., Bailer-Jones C.A.L., Herbst W. Astron. Astrophys. 430, 1005 (2005) Rotational evolution of low mass stars: The case of NGC 2264.

J/A+A/430/1049/  table1.dat table2.dat table3.dat table4.dat  (306Kb)

Jurcsik J., Sodor A., Varadi M., Szeidl B., Washuettl A , Weber M., Dekany I., Hurta Zs., Lakatos B., Posztobanyi K., Szing A., Vida K. Astron. Astrophys. 430, 1049 (2005) The Blazhko behaviour of RR Geminorum. I. CCD photometric results in 2004.

J/A+A/430/1139/  table1.dat table2.dat table3.dat  (11Kb)

Reese C., Stoecklin T., Voronin A., Rayez J.C. Astron. Astrophys. 430, 1139 (2005) Rotational excitation and de-excitation of HF molecules by He atoms.

J/A+A/430/1143/  table1.dat table2.dat table3.dat table4.dat table5.dat  (42Kb)

Bychkov V.D., Bychkova L.V., Madej J. Astron. Astrophys., 430, 1143-1154 (2005) A catalog of stellar magnetic rotational phase curves.

J/A+A/431/143/  tabla1.dat  (549Kb)

Christlieb N., Beers T.C., Thom C., Wilhelm R., Rossi S., Flynn C., Wisotzki L., Reimers D. Astron. Astrophys. 431, 143 (2005) The stellar content of the Hamburg/ESO survey. III. Field horizontal-branch stars in the Galaxy

J/A+A/431/523/  table1.dat  (5Kb)

Bernard-Salas, J., Tielens, A.G.G.M. Astron. Astrophys. 431, 523 (2005) Physical Conditions in PDRs around Planetary Nebulae.

J/A+A/431/565/  table1.dat table2.dat table3.dat  (85Kb)

Ortiz R., Lorenz-Martins S., Maciel W.J., Rangel E.M. Astron. Astrophys. 431, 565 (2005) Evolution from AGB to planetary nebula in the MSX survey.

J/A+A/431/773/  charm2.dat refs.dat  (3.7 Mb)

Richichi A., Percheron I., Khristoforova M. Astron. Astrophys. 431, 773 (2005) CHARM2: an updated Catalog of High Angular Resolution Measurements.

J/A+A/431/779/  tablea1.dat tablea2.dat  (55Kb)

Jimenez-Esteban F.M., Agudo-Merida L., Engels D., Garcia-Lario P. Astron. Astrophys. 431, 779 (2005) An infrared study of galactic OH/IR stars. I. An optical/near-IR atlas of the Arecibo sample.

J/A+A/431/793/  table2.dat table4.dat  (14Kb)

Venemans B.P., Rottgering H.J.A., Miley G.K., Kurk J.D., De Breuck C., Overzier R.A., van Breugel W.J.M., Carilli C.L., Ford H., Heckman T., Pentericci L., McCarthy P. Astron. Astrophys. 431, 793 (2005) Properties of Ly-alpha emitters around the radio galaxy MRC 0316-257.

J/A+A/431/933/  table1.dat  (11Kb)

Tautvaisiene G., Edvardsson B., Puzeras E., Ilyin I. Astron. Astrophys. 431, 933 (2005) Chemical composition of evolved stars in the open cluster NGC 7789.

J/A+A/431/1177 See VIII/74A/

J/A+A/431/1189/  lc/* stars.dat table4.dat table5.dat table6.dat  (241Kb)

Baldacci L., Rizzi L., Clementini G., Held E.V. Astron. Astrophys., 431, 1189-1201 (2005) Variable stars in the dwarf irregular galaxy NGC 6822: The photometric catalogue.

J/A+A/432/69/  table1.dat table2.dat  (19Kb)

Brusa M., Comastri A., Daddi E., Pozzetti L., Zamorani G., Vignali C., Cimatti A., Fiore F., Mignoli M., Ciliegi P., Rottgering H.J.A. Astron. Astrophys. 432, 69 (2005) XMM-Newton observations of Extremely Red Objects and the link with luminous, X-ray obscured quasars.

J/A+A/432/369/  lamda/* levels.dat molecule.dat  (6.7 Mb)

Schoeier F.L., van der Tak, F.F.S., van Dishoeck E.F., Black J.H. Astron. Astrophys. 432, 369 (2005) An atomic and molecular database for analysis of submillimetre line observations

J/A+A/432/381/  fields.dat members.dat  (21Kb)

Demarco R., Rosati P., Lidman C., Homeier N.L., Scannapieco E., Benitez N., Mainieri V., Nonino M., Girardi M., Stanford S.A., Tozzi P., Borgani S., Silk J., Squires G., Broadhurst T.J. Astron. Astrophys. 432, 381 (2005) A VLT spectroscopic survey of RX J0152.7-1357, a forming cluster of galaxies at z=0.837.

J/A+A/432/491/  table3a.dat table3b.dat  (26Kb)

Giorgi E.E., Baume G., Solivella G., Vazquez R.A. Astron. Astrophys. 432, 491 (2005) The Vela-Puppis open clusters Pismis 8 and Pismis 13.

J/A+A/432/657/  table2.dat table3.dat  (91Kb)

Jaervinen S.P., Berdyugina S.V., Tuominen I., Cutispoto G., Bos M. Astron. Astrophys. 432, 657 (2005) Magnetic activity in the young solar analog AB Dor. Active longitudes and cycles from long-term photometry.

J/A+A/432/731/  table1.dat table1.tex table2.dat table2.ps table2.tex  (324Kb)

Wilson N.J., Bell K.L., Hudson C.E. Astron. Astrophys. 432, 731 (2005) Effective collision strengths for electron impact excitation of C II.

J/A+A/432/737/  catalog.dat refs.dat  (112Kb)

Pestalozzi M.R., Minier V.

Astron. Astrophys. 432, 737 (2005) A General Catalogue of 6.7GHz Methanol Masers

J/A+A/432/851/  table34.dat  (8Kb)

Recio-Blanco A., Piotto G., De Angeli F., Cassisi S., Riello M., Salaris M., Pietrinferni A., Zoccali M., Aparicio A. Astron. Astrophys., 432, 851-859 (2005) A homogeneous set of globular cluster relative distances and reddenings.

J/A+A/432/861/  table1.dat table2.dat  (60Kb)

Gavilan M., Buell J.F., Molla M. Astron. Astrophys. 432, 861 (2005) New low and intermediate mass stars yields: The evolution of carbon abundances.

J/A+A/432/921/  tablea1.dat tablea2.dat tablea3.dat tablea4.dat tablea5.dat tablea6.dat  (62Kb)

Fontani F., Beltran M.T., Brand J., Cesaroni R., Testi L., Molinari S., Walmsley C.M. Astron. Astrophys., 432, 921-935 (2005) Search for massive protostellar candidates in the southern hemisphere. I. Association with dense gas.

J/A+A/432/1063/  table5.dat  (25Kb)

Lanzafame A.C., Brooks D.H., Lang J. Astron. Astrophys. 432, 1063 (2005) ADAS analysis of the differential emission measure structure of the inner solar corona. II. A study of the ’quiet Sun’ inhomogeneities from SOHO CDS-NIS spectra.

J/A+A/432/1137/ (3Kb) Del Zanna G., Chidichimo M.C., Mason H.E. Astron. Astrophys., 432, 1137-1150 (2005) Benchmarking atomic data for astrophysics: Fe XXIII.

J/A+A/432/1151/  table1.dat table2.dat table3.dat table4.dat  (76Kb)

Aggarwal K.M., Keenan F.P., Rose S.J. Astron. Astrophys. 432, 1151 (2005) Electron impact excitation of Al XIII : a relativistic approach.

J/A+A/433/43/  tablea1.dat tablea2.dat tablea3.dat tablea4.dat tablea5.dat tablea6.dat  (577Kb)

Dirsch B., Schuberth Y., Richtler T. Astron. Astrophys. 433, 43 (2005) A wide-field photometric study of the globular cluster system of NGC 4636.

J/A+A/433/117/  table1.dat table2.dat  (38Kb)

Viehmann T., Eckart A., Schoedel R., Moultaka J., Straubmeier C., Pott J.-U. Astron. Astrophys. 433, 117 (2005) L- and M-band imaging observations of the Galactic Center region

J/A+A/433/151/  tablea2.dat tablea3.dat tablea4.dat  (56Kb)

Zickgraf F.-J., Krautter J., Reffert S., Alcala J.M., Mujica R., Covino E., Sterzik M.F. Astron. Astrophys., 433, 151-171 (2005) Identification of a complete sample of northern ROSAT All-Sky Survey X-ray sources. VIII. The late-type stellar component.

J/A+A/433/185/  table4.dat table8.dat table9.dat  (53Kb)

Bensby T., Feltzing S., Lundstroem I., Ilyin I. Astron. Astrophys. 433, 185 (2005) α-, r-, and s-process element abundance trends in the Galactic thin and thick disks.

J/A+A/433/497/  table6.dat table7.dat table9.dat table10.dat  (231Kb)

Rampazzo R., Annibali F., Bressan A., Longhetti M., Padoan F., Zeilinger W.W. Astron. Astrophys. 433, 497 (2005) Nearby early-type galaxies with ionized gas. I. Line-strength indices of the underlying stellar population.

J/A+A/433/731/  table2.dat table3.dat  (487Kb)

Del Zanna G., Mason H.E. Astron. Astrophys. 433, 731 (2005) Benchmarking atomic data for astrophysics: Fe XII.

J/A+A/433/745/  table1.dat table8.dat table9.dat table10.dat  (3.0 Mb)

Jonauskas V., Bogdanovich P., Keenan F.P., Foord M. E., Heeter R.F., Rose S.J., Ferland G.J., Kisielius R., van Hoof P.A.M., Norrington P.H. Astron. Astrophys. 433, 745 (2005) Energy levels and transition probabilities for nitrogen-like Fe XX.

J/A+A/433/751/  table1.dat table2.dat  (12Kb)

Makarova L.N., Karachentsev I.D., Grebel E.K., Harbeck D., Korotkova G.G., Geisler D. Astron. Astrophys., 433, 751-756 (2005) Imaging and photometry of nearby dwarf galaxies. II. Southern dwarfs.

J/A+A/433/757/  table4.dat  (26Kb)

Sluse D., Hutsemekers D., Lamy H., Cabanac R., Quintana H. Astron. Astrophys. 433, 757 (2005) New optical polarization measurements of quasi-stellar objects. The data.

J/A+A/433/1155/  calib.dat  (132Kb)

Merand A., Borde P., Coude du Foresto V. Astron. Astrophys. 433, 1155 (2005) A catalog of bright calibrator stars for 200-meter baseline near-infrared stellar interferometry.

J/A+A/433/1163/  table1.dat table2.dat table3.dat table4.dat table5.dat  (59Kb)

Donato D., Sambruna R.M., Gliozzi M. Astron. Astrophys., 433, 1163-1169 (2005) Six years of BeppoSAX observations of blazars: A spectral catalog.

J/A+A/434/53/  table2.dat table2.fit  (70Kb)

Vanzella E., Cristiani S., Dickinson M., Kuntschner H., Moustakas L.A., Nonino M., Rosati P., Stern D., Cesarsky C., Ettori S., Ferguson H.C., Fosbury R.A.E., Giavalisco M., Haase J., Renzini A., Rettura A., Serra P. Astron. Astrophys. 434, 53 (2005) The Great Observatories Origins Deep Survey VLT/FORS2 Spectroscopy in the GOODS-South Field.

J/A+A/434/89/  table1.dat table3.dat  (34Kb)

Michalska G., Pigulski A. Astron. Astrophys. 434, 89 (2005) Detached binaries in the Large Magellanic Cloud. A selection of binaries suitable for distance determination.

J/A+A/434/149/  galaxies.dat table3.dat  (10Kb)

Valdes J.R., Berta S., Bressan A., Franceschini A., Rigopoulou D., Rodighiero G. Astron. Astrophys. 434, 149 (2005) NIR spectroscopy of luminous infrared galaxies and the hydrogen recombination photon deficit.

J/A+A/434/201/  refs.dat table1.dat table1a.dat  (34Kb)

He J.H., Szczerba R., Chen P.S., Sobolev A.M. Astron. Astrophys. 434, 201 (2005) A search for radiative pumping lines of OH masers.

J/A+A/434/235/  ews.dat stars.dat  (29Kb)

Fabbian D., Recio-Blanco A., Gratton R.G., Piotto G. Astron. Astrophys. 434, 235 (2005) Abundance anomalies in hot horizontal branch stars of the Galactic globular cluster NGC 1904.

J/A+A/434/275/  stars.dat table6.dat  (161Kb)

del Peloso E.F., da Silva L., Porto de Mello G.F. Astron. Astrophys. 434, 275 (2005) The age of the Galactic thin disk from Th/Eu nucleocosmochronology. I. Determination of [Th/Eu] abundance ratios.

J/A+A/434/365/  aos.dat table1.dat  (46Kb)

Landi O. Astron. Astrophys. 434, 365 (2005) Radiative transition probabilities in the O-like sequence.

J/A+A/434/385/  refs.dat table1.dat table2.dat table3.dat  (25Kb)

Giommi P., Piranomonte S., Perri M., Padovani P. Astron. Astrophys. 434, 385 (2005) The Sedentary Survey of Extreme High Energy Peaked BL Lacs. II. The Catalog and Spectral Properties.

J/A+A/434/475/  table1.dat table2.dat  (18Kb)

Kitsionas S., Hatziminaoglou E., Georgakakis A., Georgantopoulos I. Astron. Astrophys. 434, 475 (2005) On the use of photometric redshifts for X-ray selected AGNs.

J/A+A/434/483/  refs.dat table1.dat table2.dat  (628Kb)

Pietsch W., Freyberg M., Haberl F. Astron. Astrophys. 434, 483 (2005)

An XMM-Newton survey of M 31.

J/A+A/434/613/  table1.dat table2.dat table3.dat  (15Kb)

Szymczak M., Pillai T., Menten K.M. Astron. Astrophys. 434, 613 (2005) Masers as signposts of high-mass protostars. A water maser survey of methanol maser sources.

J/A+A/434/773/  table3.dat  (12Kb)

Mahmoudi W.F., Ben Nessib N., Dimitrijevic M.S. Astron. Astrophys. 434, 773 (2005) Modified semiempirical electron width calculations of singly-ionized oxygen spectral lines.

J/A+A/434/801/  catalog.dat  (9Kb)

Zappacosta L., Maiolino R., Finoguenov A., Mannucci F., Gilli R., Ferrara A. Astron. Astrophys., 434, 801-809 (2005) Constraining the thermal history of the warm-hot intergalactic medium.

J/A+A/434/887/  table1.dat table2.dat  (17Kb)

Huchtmeier W.K., Gopal Krishna, Petrosian A. Astron. Astrophys. 434, 887 (2005) HI-observations of blue compact dwarf galaxies.

J/A+A/434/987/  table5.dat  (20Kb)

Apai D., Linz H., Henning T., Stecklum B. Astron. Astrophys. 434, 987 (2005) Infrared portrait of of the nearby massive star-forming region IRAS 09002-4732.

J/A+A/434/1107/  photom.dat table3.dat table4.dat  (226Kb)

Munari U., Henden A., Vallenari A., Bond H.E., Corradi R.L.M., Crause L., Desidera S., Giro E., Marrese P.M., Ragaini S., Siviero A., Sordo R., Starrfield S., Tomov T., Villanova S., Zwitter T., Wagner R.M. Astron. Astrophys. 434, 1107 (2005) On the distance, reddening and progenitor of V838 Mon.

J/A+A/434/1125/  table1.dat table4.dat  (21Kb)

Dufton P.L., Ryans R.S.I., Trundle, C., Lennon D.J., Hubeny I., Lanz T., Allende Prieto C. Astron. Astrophys. 434, 1125 (2005) B-type Supergiants in the SMC: Chemical Compositions and comparison of static and unified models.

J/A+A/434/1201/  table4.dat  (49Kb)

Richichi A., Percheron I. Astron. Astrophys. 434, 1201 (2005) First results from the ESO VLTI Calibrators Program.

J/A+A/435/65/  table2.dat  (20Kb)

Bastian N., Hempel H., Kissler-Patig M., Homeier N., Trancho G. Astron. Astrophys. 435, 65 (2005) The star cluster population in the tidal tails of NGC 6872.

J/A+A/435/363/  table5.dat table6.dat  (26Kb)

Deacon N.R., Hambly N.C., Cooke J.A. Astron. Astrophys. 435, 363 (2005) Southern Infrared Proper Motion Survey. I. Discovery of New High Proper Motion Stars From First Full Hemisphere Scan.

J/A+A/435/373/  table1.dat table2.dat  (2.4 Mb)

Barklem P.S., Aspelund-Johansson J. Astron. Astrophys. 435, 373 (2005) The broadening of Fe II lines by neutral hydrogen collisions.

J/A+A/435/437/  tablea1.dat  (14Kb)

Urosevic D., Pannuti T.G., Duric N., Theodorou A. Astron. Astrophys., 435, 437-447 (2005) The Σ-D relation for supernova remnants in nearby galaxies.

J/A+A/435/459/  table1.dat table2.dat  (27Kb)

Huchtmeier W.K., Karachentsev I.D., Karachentseva V.E., Kudrya Yu.N., Mitronova S.N. Astron. Astrophys. 435, 459 (2005) HI observations of edge-on spiral galaxies.

J/A+A/435/507/  tablea1.dat  (15Kb)

Zheng X.Z., Hammer F., Flores H., Assemat F., Rawat A. Astron. Astrophys. 435, 507 (2005)

HST/WFPC2 morphologies and bar structures of field galaxies at 0.4 z 1.

J/A+A/435/521/  notes.dat table1.dat table2.dat  (42Kb)

Nagar N.M., Falcke H., Wilson A.S. Astron. Astrophys. 435, 521 (2005) Radio Sources in Low-Luminosity Active Galactic Nuclei. IV. Radio Luminosity Function, Importance of Jet Power, and Radio Properties of the Complete Palomar Sample

J/A+A/435/575/  table3.dat  (16Kb)

Messineo M., Habing H.J., Menten K.M., Omont A., Sjouwerman L.O., Bertoldi F. Astron. Astrophys. 435, 575 (2005) 86 GHz SiO maser survey of late-type stars in the Inner Galaxy. III. Interstellar extinction and colours of the SiO targets

J/A+A/435/657/  table1.dat tablea1.dat  (52Kb)

Alves-Brito A., Barbuy B., Ortolani S., Momany I., Hill V., Zoccali M., Renzini A., Minniti D., Pasquini L., Bica E., Rich R.M. Astron. Astrophys. 435, 657 (2005) VLT-UVES analysis of 5 giants in 47 Tucanae.

J/A+A/435/781/  table3.dat table5.dat  (46Kb)

Olsen L.F., Benoist C., da Costa L., Hansen L., Joergensen H.E. Astron. Astrophys. 435, 781 (2005) Multi-object spectroscopy of low redshift EIS clusters. III. Properties of optically selected clusters.

J/A+A/435/799/  table1.dat  (31Kb)

Tajer M., Trinchieri G., Wolter A., Campana S., Moretti A., Tagliaferri G. Astron. Astrophys., 435, 799-810 (2005) A sample of X-ray emitting normal galaxies from the BMW-HRI Catalogue.

J/A+A/435/863/  table2.dat table3.dat  (20Kb)

Mack K.-H., Vigotti M., Gregorini L., Klein U., Tschager W., Schilizzi R.T., Snellen I.A.G. Astron. Astrophys., 435, 863-870 (2005) Multi-frequency study of the B3VLA sample. IV. 74-MHz flux densities from VLA A-array data.

J/A+A/435/1087/  table2.dat table3.dat table4.dat table5.dat table6.dat  (40Kb)

Luna G.J.M., Costa R.D.D. Astron. Astrophys., 435, 1087-1097 (2005) Nebular abundances of southern symbiotic stars.

J/A+A/435/1099/  table2.dat table3.dat table4.dat table5.dat tablea.dat  (30Kb)

Adelman S.J., Kaewkornmaugh P. Astron. Astrophys. 435, 1099 (2005) FCAPT photometry of the mCP stars HD 32633, theta Aur, 49 Cam, and 3 Hya.

J/A+A/436/91/  table2.dat  (40Kb)

Demers S., Battinelli P. Astron. Astrophys. 436, 91 (2005) The C star outer disk population of M31 seen with the SLOAN filters.

J/A+A/436/117/  table2.dat table3.dat  (162Kb)

Meynadier F., Heydari-Malayeri M., Walborn N.R. Astron. Astrophys. 436, 117 (2005) The LMC HII region N 214C and its peculiar nebular blob.

J/A+A/436/443/  refs.dat table1.dat table2.dat table3.dat  (71Kb)

Verdes-Montenegro L., Sulentic J., Lisenfeld U., Leon S., Espada D., Garcia E., Sabater J., Verley S. Astron. Astrophys. 436, 443 (2005) The AMIGA project. I. Optical characterization of the Catalogue of Isolated Galaxies.

J/A+A/436/457/  notes.dat table1.dat  (32Kb)

Gopal-Krishna, Ledoux C., Melnick J., Giraud E., Kulkarni V., Altieri B. Astron. Astrophys., 436, 457-464 (2005) Optical and radio observations of a sample of 52 powerful ultra-steep spectrum radio sources.

J/A+A/436/527/  table5.dat  (16.1 Mb)

Carraro G., Baume G., Piotto G., Mendez R.A., Schrnidtobreik L. Astron. Astrophys. 436, 527 (2005) Luminosity and mass functions of galactic open clusters: II. NGC 4852.

J/A+A/436/687/  table1.dat  (8Kb)

Hunter I., Dufton P.L., Ryans R.S.I., Lennon D.J., Rolleston W.R.J., Hubeny I., Lanz T.

Astron. Astrophys., 436, 687-695 (2005) A non-LTE analysis of the spectra of two narrow lined main sequence stars in the SMC.

J/A+A/436/799/  refs.dat table1.dat  (9Kb)

Fan J.H. Astron. Astrophys., 436, 799-804 (2005) The basic parameters of γ-ray-loud blazars.

J/A+A/436/853/  table2.dat  (2.5 Mb)

Lodieu N., McCaughrean M.J., Barrado y Navascues D., Bouvier J., Stauffer J.R. Astron. Astrophys. 436, 853 (2005) A near-infrared survey for new low-mass members in Alpha Per.

J/A+A/436/1131/  j-index.dat j-index.tex table7.dat table7a.tex table7b.tex  (72Kb)

Hudson C.E., Bell K.L. Astron. Astrophys. 436, 1131 (2005) Fine structure effective collision strengths for the electron impact excitation of Al III.

J/A+A/436/1141/ table1.dat table2.dat table3.dat table4.dat table5.dat table6.dat table7.dat table8.dat table9.dat table10.dat table11.dat tables.ps tables.tex (125.7 Mb) Aggarwal K.M., Keenan F.P., Nakazaki S. Astron. Astrophys. 436, 1141 (2005) Energy levels and radiative rates for transitions in Ar XIII, Ar XIV and Ar XV.

J/A+A/437/61/  table2.dat  (13Kb)

Kang A., Sohn Y.-J., Rhee J., Shin M., Chun M.-S., Kim H.-I. Astron. Astrophys. 437, 61 (2005) Near-IR photometry of asymptotic giant branch stars in the dwarf elliptical galaxy NGC 185.

J/A+A/437/87/ (2Kb) Doroshenko V.T., Sergeev S.G., Merkulova N.I., Sergeeva E.A., Golubinsky Y.V. Astron. Astrophys., 437, 87-93 (2005) Optical continuum properties of the NLS1 galaxy Mrk 335 in 1995-2004.

J/A+A/437/101/  table1.dat  (54Kb)

Ehlerova S., Palous J. Astron. Astrophys., 437, 101-112 (2005) HI shells in the outer Milky Way.

J/A+A/437/345/  enid-fe4.user fjj-fe4.user fls-fe4.user lifetime.f table3.dat table4.dat  (100.4 Mb)

Nahar S.N., Pradhan A.K. Astron. Astrophys. 437, 345 (2005) Atomic data from the Iron Project. LIX. New radiative transition probabilities for Fe IV including fine structure.

J/A+A/437/375/  catalog.dat refs.dat  (20Kb)

Dryomova G.N., Perevozkina E.L., Svechnikov M.A. Astron. Astrophys. 437, 375 (2005) Catalog of orbital elements, masses and luminosities of short-periodic RS CVn-type eclipsing systems.

J/A+A/437/457/  table3.dat table4.dat table5.dat  (1.0 Mb)

Balaguer-Nunez L., Jordi C., Galadi-Enriquez D. Astron. Astrophys. 437, 457 (2005) uvby-HβCCD photometry and membership segregation of the open cluster NGC 2548; Gaps in the Main Sequence of open clusters.

J/A+A/437/467/  table9.dat table10.dat table11.dat table12.dat  (47Kb)

Evans C.J., Smartt S.J., Lee J.-K., Lennon D.J., Kaufer A., Dufton P.L., Trundle C., Herrero A., Simon-diaz S., De Koter A., Hamann W.-R., Hendry M.A., Hunter I., Irwin M.J., Korn A.J., Kudritzki R.-P., Langer N., Mokiem M.R., Najarro F., Pauldrach A.W.A., Przybilla N., Puls J., Ryans R.S.I., Urbaneja M.A., Venn K.A., Villamariz M.R. Astron. Astrophys., 437, 467-482 (2005) The VLT-FLAMES survey of massive stars: Observations in the Galactic clusters NGC 3293, NGC 4755 and NGC 6611.

J/A+A/437/705/ 7a7.dat 7ai.dat 7bp.dat 79j.dat 79l.dat 79m.dat 79p1.dat 79p2.dat 79t1.dat d2i.dat d2j1.dat d2j2.dat d2r.dat d3f.dat d3i.dat d3o.dat d3q.dat d4b.dat d18.dat d23.dat (13.0 Mb) Pauwels T., Vingerhoets P., Cuypers J., Lampens P. Astron. Astrophys. 437, 705 (2005) Photometric observations of the mutual phenomena of the Galilean Satellites of Jupiter in 1997 and 2003 at the Royal Observatory of Belgium.

J/A+A/437/789/  table2.dat  (10Kb)

Nobili S., Amanullah R., Garavini G., Goobar A., Lidman C., Stanishev V., Aldering G., Antilogus P., Astier P., Burns M.S., Conley A., Deustua S.E., Ellis R., Fabbro S., Fadeyev V., Folatelli G., Gibbons R., Goldhaber G., Groom D.E., Hook I., Howell D.A., Kim A.G., Knop R.A., Nugent P.E., Pain R., Perlmutter S., Quimby R., Raux J., Regnault N., Ruiz-Lapuente P., Sainton G., Schahmaneche K., Smith E., Spadafora A.L., Thomas R.C., Wang L., (the Supernova Cosmology Project) Astron. Astrophys., 437, 789-804 (2005) Restframe I-band Hubble diagram for type Ia supernovae up to redshift z 0.5.

J/A+A/437/883/  catalog.dat sp/*  (24.6 Mb)

Mignoli M., Cimatti A., Zamorani G., Pozzetti L., Daddi E., Renzini A., Broadhurst T., Cristiani S., D’Odorico S., Fontana A., Giallongo E., Gilmozzi R., Menci N., Saracco P. Astron. Astrophys. 437, 883 (2005) The K20 survey - VII. The spectroscopic catalogue: Spectral properties and evolution of the galaxy population.

J/A+A/437/1135/  position.dat table2.dat table3.dat table4.dat  (24Kb)

Camargo I.B., Daigne G., Ducourant C., Charlot P. Astron. Astrophys. 437, 1135 (2005) Near-infrared astrometry and photometry of Southern ICRF quasars.

J/A+A/437/1147/  table2.dat  (34Kb)

Veiga C.H., Vieira Martins R. Astron. Astrophys., 437, 1147-1150 (2005) CCD astrometric observations of Amalthea and Thebe.

J/A+A/438/139/  catalog.dat param.dat  (143Kb)

Soubiran C., Girard P. Astron. Astrophys. 438, 139 (2005) Abundance trends in kinematical groups of the Milky Way’s disk.

J/A+A/438/201/  dataset1.dat dataset2.dat  (157Kb)

Lampens P., Fremat Y., Garrido R., Pena J.H., Parrao L., Van Cauteren P., Cuypers J., De Cat P., Uytterhoeven K., Arentoft T., Hobart M. Astron. Astrophys. 438, 201 (2005) A photometric study of the light variations of the triple system DG Leonis.

J/A+A/438/251/  table6.dat table45.dat  (15Kb)

Beirao P., Santos N.C., Israelian G., Mayor M. Astron. Astrophys., 438, 251-256 (2005) Abundances of Na, Mg and Al in stars with giant planets.

J/A+A/438/291/  table1.dat  (10Kb)

Gondoin P. Astron. Astrophys., 438, 291-300 (2005) X-ray observations of the old open stellar cluster NGC 188.

J/A+A/438/377/  kmac1cu.dat kmac1t.dat list.dat  (32.4 Mb)

Lazorenko P., Babenko Yu, Karbovsky V., Buromsky M., Denisjuk O., Kasjan S. Astron. Astrophys. 438, 377 (2005) The Kyiv Meridian Axial Circle Catalogue of stars in fields with extragalactic radio sources.

J/A+A/438/507/  tablea1.dat  (11Kb)

Bournaud F., Combes F., Jog C.J., Puerari I. Astron. Astrophys., 438, 507-520 (2005) Lopsided spiral galaxies: evidence for gas accretion.

J/A+A/438/521/  table1.dat table3.dat  (290Kb)

Raimondo G., Cioni M.-R.L., Rejkuba M., Silva D.R. Astron. Astrophys., 438, 521-532 (2005) Pulsation properties of C stars in the Small Magellanic Cloud.

J/A+A/438/663/  table6.dat  (21Kb)

Maercker M., Burton M.G. Astron. Astrophys. 438, 663 (2005) L-band (3.5 micron) IR-excess in massive star formation. I. 30 Doradus.

J/A+A/438/769/  pms.dat  (202Kb)

Ducourant C., Teixeira R., Perie J.P., Lecampion J.F., Guibert J., Sartori, M.J. Astron. Astrophys. 438, 769 (2005) Pre-main sequence stars Proper Motion Catalogue.

J/A+A/438/1163/  atlas/* cluster.dat cocd.dat notes.dat  (16.4 Mb)

Kharchenko N.V., Piskunov A.E., Roeser S., Schilbach E., Scholz R.-D. Astron. Astrophys. 438, 1163 (2005) Astrophysical parameters of Galactic open clusters.

J/A+A/438/1175/  table2.dat  (39Kb)

Rau A., Kienlin A. von, Hurley K., Lichti G.G. Astron. Astrophys. 438, 1175 (2005) The 1st INTEGRAL SPI-ACS Gamma-Ray Burst Catalogue.

J/A+A/439/129/  table1.dat table2a.dat table2b.dat  (337Kb)

Barklem P.S., Christlieb N., Beers T.C., Hill V., Bessell M.S., Holmberg J., Marsteller B., Rossi S., Zickgraf F.-J., Reimers D. Astron. Astrophys., 439, 129-151 (2005) The Hamburg/ESO R-process enhanced star survey (HERES). II. Spectroscopic analysis of the survey sample.

J/A+A/439/265/  table1.dat tablea1.dat tablea2.dat  (44Kb)

Hadfield L.J., Crowther P.A., Schild H., Schmutz W. Astron. Astrophys., 439, 265-277 (2005) A spectroscopic search for the non-nuclear Wolf-Rayet population of the metal-rich spiral galaxy M83.

J/A+A/439/387/  fiv.dat fv.dat nev.dat niii.dat niv.dat oiv.dat ov.dat  (187Kb)

Bizau J.-M., Champeaux J.-P., Cubaynes D., Wuilleumier F.J., Folkmann F., Jacobsen T.S., Penent F., Blancard C., Kjeldsen H. Astron. Astrophys. 439, 387 (2005) Absolute cross sections for L-shell photoionization of the ions N2+, N3+, O3+, O4+, F3+, F4+ and Ne4+.

J/A+A/439/413/  xmds1.dat  (65Kb)

Chiappetti L., Tajer M., Trinchieri G., Maccagni D., Maraschi L., Paioro L., Pierre M., Surdej J., Garcet O., Gosset E., Le Fevre O., Bertin E., McCracken H.J., Mellier Y., Foucaud S., Radovich M., Ripepi V., Arnaboldi M. Astron. Astrophys. 439, 413 (2005) The XMM-LSS survey: The XMDS/VVDS 4σcatalogue.

J/A+A/439/433/  table2.dat  (234Kb)

Janssen T., Van Kerkwijk M.H. Astron. Astrophys. 439, 433 (2005) Observations of the binary pulsar system PSR B1718-19. The role of tidal circularization.

J/A+A/439/487/  table1.dat table2.dat table3.dat table4.dat  (16Kb)

De Ruiter H.R., Parma P., Capetti A., Fanti R., Morganti R., Santantonio L. Astron. Astrophys., 439, 487-496 (2005) Are radio galaxies and quiescent galaxies different? Results from the analysis of HST brightness profiles.

J/A+A/439/559/  table2.dat tableb1.dat  (192Kb)

Groenewegen M.A.T. Astron. Astrophys. 439, 559 (2005) Eclipsing binaries in the Galactic Bulge: candidates for distance estimates.

J/A+A/439/845 See III/250/

J/A+A/439/1107/  atlas.dat sp/*  (266Kb)

Davies B., Oudmaijer R.D., Vink J.S. Astron. Astrophys. 439, 1107 (2005) Asphericity and Clumpiness in the winds of Luminous Blue Variables.

J/A+A/439/1137/  stars.dat table1.dat  (193Kb)

Fuhrmeister B., Schmitt J.H.M.M., Hauschildt P.H. Astron. Astrophys. 439, 1137 (2005) PHOENIX model chromospheres of mid- to late-type M dwarfs.

J/A+A/439/1149/  tablea1.dat  (10Kb)

Argiroffi C., Maggio A., Peres G., Stelzer B., Neuhaeuser R. Astron. Astrophys., 439, 1149-1158 (2005) XMM-Newton spectroscopy of the metal depleted T Tauri star TWA 5.

J/A+A/439/1215/  table1.dat table2.dat table5.dat table6.dat tables.tex  (1.2 Mb)

Aggarwal K.M., Keenan F.P. Astron. Astrophys. 439, 1215 (2005) Effective collision strengths for transitions in Fe X.

J/A+A/440/53/  table1.dat table2.dat  (15Kb)

Gaynullina E.R., Schmidt R.W., Akhunov T. Burkhonov O., Gottloeber S., Mirtadjieva K., Nuritdinov S.N., Tadjibaev I., Wambsganss J., Wisotzki L. Astron. Astrophys. 440, 53 (2005) Microlensing in the double quasar SBS1520+530.

J/A+A/440/61/  table2.dat table3.dat  (24Kb)

Rigopoulou D., Vacca W.D., Berta S., Franceschini A., Aussel H. Astron. Astrophys. 440, 61 (2005) FORS spectroscopy of galaxies in the Hubble Deep Field South.

J/A+A/440/107/  table1.dat  (20Kb)

Coccato L., Corsini E.M., Pizzella A., Bertola F. Astron. Astrophys. 440, 107 (2005) Spiral galaxies with a central plateau in the gas velocity curve along the major axis.

J/A+A/440/111/  clusters.dat spec/*  (541Kb)

Piatti A.E., Santos J.F.C.Jr, Claria J.J., Bica E., Ahumada A.V., Parisi M.C. Astron. Astrophys. 440, 111 (2005) Integrated spectral analysis of 18 concentrated star clusters in the Small Magellanic Cloud.

J/A+A/440/151/  refs.dat tablea1.dat  (15Kb)

Hatchell J., Richer J.S., Fuller G.A., Qualtrough C.J., Ladd E.F., Chandler C.J. Astron. Astrophys., 440, 151-161 (2005) Star formation in Perseus. Clusters, filaments and the conditions for star formation.

J/A+A/440/249/  table1.dat  (12Kb)

Uytterhoeven K., Harmanec P., Telting J.H., Aerts C. Astron. Astrophys. 440, 249 (2005) The orbit of the close spectroscopic binary epsilon Lupi and the intrinsic variability of its early B-type components

J/A+A/440/305/  table9.dat  (16Kb)

Fremat Y., Zorec J., Hubert A.-M., Floquet M. Astron. Astrophys., 440, 305-320 (2005) Effects of gravitational darkening on the determination of fundamental parameters in fast-rotating B-type stars.

J/A+A/440/321/  table1.dat table2.dat table3.dat table4.dat  (79Kb)

Jonsell K., Edvardsson B., Gustafsson B., Magain P., Nissen P.E., Asplund M. Astron. Astrophys. 440, 321 (2005) Chemical abundances in 43 metal-poor stars.

J/A+A/440/373/  table1.dat  (163Kb)

Howard T.A., Tappin S.J. Astron. Astrophys. 440, 373 (2005) Statistical survey of earthbound interplanetary shocks, associated coronal mass ejections and their space weather consequences.

J/A+A/440/403/  atlas/* clusters.dat index.dat notes.dat stars.dat  (9.6 Mb)

Kharchenko N.V., Piskunov A.E., Roeser S., Schilbach E., Scholz R.-D. Astron. Astrophys., 440, 403-408 (2005) 109 new Galactic open clusters.

J/A+A/440/409/  position.dat table2.dat  (535Kb)

Terasranta H., Wiren S., Koivisto P., Saarinen V., Hovatta T. Astron. Astrophys. 440, 409 (2005) 24 year monitoring of extragalactic sources at 22 and 37 GHz.

J/A+A/440/487/  table1.dat table2.dat table4.dat  (32Kb)

Storm J., Gieren W.P., Barnes III T.G., Gomez, M. Astron. Astrophys. 440, 487 (2005) The near-IR Surface Brightness Method applied to six Cepheids in the young LMC cluster NGC 1866.

J/A+A/440/595/  table1.dat  (7Kb)

Alencar S.H.P., Basri G., Hartmann L., Calvet N. Astron. Astrophys., 440, 595-608 (2005) The extreme T Tauri star RW Aur: accretion and outflow variability.

J/A+A/440/647/  grid.dat x748z004/* x754z002/* z002.dat z004.dat  (62.1 Mb)

Claret A. Astron. Astrophys. 440, 647 (2005)

New grids of stellar models including tidal-evolution constants up to carbon burning. II. From 0.8 to 125M⊙: the Small Magellanic Cloud (Z=0.002-0.004).

J/A+A/440/676 See VIII/75/

J/A+A/440/735/  table2.dat  (115Kb)

Jaervinen S.P., Berdyugina S.V., Strassmeier K.G. Astron. Astrophys. 440, 735 (2005) Spots on EK Draconis - Active longitudes and cycles from long-term photometry.

J/A+A/440/775 See VIII/76/

J/A+A/440/783/  table1.dat  (11Kb)

Pietrzynski G., Ulaczyk K., Gieren W., Bresolin F., Kudritzki R.P. Astron. Astrophys., 440, 783-789 (2005) A survey for OB associations in the Sculptor Group spiral galaxy NGC 7793.

J/A+A/440/901/  table3.dat table4.dat table7.dat  (25Kb)

Gratton R.G., Bragaglia A., Carretta E., De Angeli F., Lucatello S., Momany Y., Piotto G., Recio Blanco A. Astron. Astrophys. 440, 901 (2005) Precise reddening and metallicity of NGC 6752 from FLAMES spectra.

J/A+A/440/1061/ main.dat obslog1.dat obslog2.dat obslog3.dat table5.dat table6.dat table7.dat (43Kb) Lodieu N., Scholz R.-D., Mccaughrean M.J., Ibata R., Irwin M., Zinnecker H. Astron. Astrophys., 440, 1061-1078 (2005) Spectroscopic classification of red high proper motion objects in the Southern Sky.

J/A+A/440/1153/  table5.dat  (17Kb)

del Peloso E.F., da Silva L., Arany-Prado L.I. Astron. Astrophys. 440, 1153 (2005) The age of the Galactic thin disk from Th/Eu nucleocosmochronology. III. Extended sample.

J/A+A/441/181/ (3Kb) Casassus S., Stahl O., Wilson T.L. Astron. Astrophys., 441, 181 (2005) Interstellar ^{12}C/^{13}C ratios through CH^{+}λλ3957,4232 absorption in local clouds: incomplete mixing in the ISM.

J/A+A/441/195/  table3.dat  (16Kb)

Vicente S.M., Alves J. Astron. Astrophys., 441, 195 (2005) Size distribution of circumstellar disks in the Trapezium cluster.

J/A+A/441/391/  table1.dat table2.dat  (140Kb)

Simic Z., Dimitrijevic M.S., Milovanovic N., Sahal-Brechot S. Astron. Astrophys. 441, 391 (2005) Stark broadening of Cd I spectral lines.

J/A+A/441/451/  photom.dat  (15Kb)

Michard R. Astron. Astrophys., 441, 451-464 (2005) Near infra-red and optical colour gradients in E-type galaxies. Inferences on dust content.

J/A+A/441/631/  table4.dat  (12Kb)

Paunzen E., Stuetz C., Maitzen H.M. Astron. Astrophys., 441, 631-640 (2005) On the detection of chemically peculiar stars using Δa photometry.

J/A+A/441/653/  table1.dat table3.dat  (17Kb)

Crifo F., Phan-Bao N., Delfosse X., Forveille T., Guibert J., Martin E.L., Reyle C. Astron. Astrophys., 441, 653-661 (2005) New neighbours. VI. Spectroscopy of DENIS nearby stars candidates.

J/A+A/441/685/  table2.dat  (6Kb)

Val Baker A., Norton A.J., Quaintrell H. Astron. Astrophys. 441, 685 (2005) The mass of the neutron star in SMC X-1.

J/A+A/441/831/ table1.dat table2.dat table3.dat table4.dat table5.dat table6.dat table7.dat table8.dat (352Kb) Aggarwal K.M., Keenan F.P. Astron. Astrophys. 441, 831 (2005) Electron impact excitation of Ar XVII.

J/A+A/441/905/  tablea1.dat tablea2.dat  (110Kb)

Hildebrandt H., Bomans D.J., Erben T., Schneider P., Schirmer M., Czoske O., Dietrich J.P., Schrabback T., Simon P., Dettmar R.J., Haberzettl L., Hetterscheidt M., Cordes O. Astron. Astrophys. 441, 905 (2005) GaBoDS: The Garching-Bonn Deep Survey. III. Lyman-Break Galaxies in the Chandra Deep Field South.

J/A+A/441/915/  position.dat refs.dat tablea1.dat  (42Kb)

Hutsemekers D., Cabanac R., Lamy H., Sluse D. Astron. Astrophys. 441, 915 (2005) Mapping extreme-scale alignments of quasar polarization vectors

J/A+A/441/981/  ngc1232.dat ngc1365.dat ngc2903.dat ngc2997.dat ngc5236.dat  (26Kb)

Bresolin F., Schaerer D., Gonzalez Delgado R.M., Stasinska G. Astron. Astrophys. 441, 981 (2005) A VLT study of metal-rich extragalactic H II regions. I. Observations and empirical abundances.

J/A+A/441/1149/  stars.dat table2.dat table3.dat  (21Kb)

del Peloso E.F., Cunha K., da Silva L., Porto de Mello G.F. Astron. Astrophys. 441, 1149 (2005) Abundances of Mn, Co and Eu in a sample of 20 F-G disk stars: the influence of hyperfine structure splitting.

J/A+A/441/1195/  ACAP/* NCAP/* PCAP/* PHOTO/*  (55.2 Mb)

Aikawa M., Arnould M., Goriely S., Jorissen A., Takahashi K. Astron. Astrophys. 441, 1195 (2005) BRUSLIB and NETGEN: the Brussels nuclear reaction rate library and nuclear network generator for astrophysics.

J/A+A/442/85/  table1.dat table2.dat table3.dat  (33Kb)

Sharina M.E., Puzia T.H., Makarov D.I. Astron. Astrophys., 442, 85-95 (2005) Hubble Space Telescope imaging of globular cluster candidates in low surface brightness dwarf galaxies.

J/A+A/442/109/  table3.dat table4.dat table5.dat table6.dat table7.dat  (43Kb)

Pustilnik S.A., Engels D., Lipovetsky V.A., Kniazev A.Y., Pramskij A.G., Ugryumov A.V., Masegosa J., Izotov Y.I., Chaffee F., Marquez I., Teplyakova A.L., Hopp U., Brosch N., Hagen H.-J., Martin J.-M. Astron. Astrophys. 442, 109 (2005) The Hamburg/SAO Survey for emission-line galaxies. VI. The Sixth List of 126 Galaxies.

J/A+A/442/137/  previous.dat table1.dat table2.dat table3.dat tables.tex  (68Kb)

Noordermeer E., Van Der Hulst J.M., Sancisi R., Swaters R.A., Van Albada T.S. Astron. Astrophys., 442, 137-157 (2005) The Westerbork HI survey of spiral and irregular galaxies. III. HI observations of early-type disk galaxies.

J/A+A/442/165/  table1.dat  (169Kb)

Cioni M.-R.L., Habing H.J. Astron. Astrophys. 442, 165 (2005) The Draco dwarf galaxy in the near-infrared.

J/A+A/442/211/  notes.dat refs.dat table1.dat  (48Kb)

Scholz R.-D., Meusinger H., Jahreiss H. Astron. Astrophys., 442, 211-227 (2005) Search for nearby stars among proper motion stars selected by optical-to-infrared photometry. III. Spectroscopic distances of 322 NLTT stars.

J/A+A/442/229/  table1.dat  (24Kb)

Maintz G., De Boer K.S. Astron. Astrophys., 442, 229-237 (2005) RR Lyrae stars: kinematics, orbits and z-distribution.

J/A+A/442/249/  table2.dat table2.tex table4.dat table4.tex  (225Kb)

Zhang Y., Liu X.-W., Luo S.-G., Pequignot D., Barlow M.J. Astron. Astrophys. 442, 249 (2005) Integrated spectrum of the planetary nebula NGC 7027.

J/A+A/442/281/  table2.dat  (110Kb)

Kucinskas A., Hauschildt P.H., Ludwig H.-G., Brott I., Vansevicius V., Lindegren L., Tanabe T., Allard F. Astron. Astrophys. 442, 281 (2005) Broad-band photometric colors and effective temperature calibrations for late-type giants. I. Z=0.02.

J/A+A/442/365/  refs.dat table1.dat table3.dat table4.dat table5.dat table6.dat  (30Kb)

Jancart S., Jorissen A., Babusiaux C., Pourbaix D.

Astron. Astrophys., 442, 365-380 (2005) Astrometric orbits of SB9 stars.

J/A+A/442/381/  notes.dat refs.dat table1.dat  (72Kb)

Maintz G. Astron. Astrophys. 442, 381 (2005) Proper identification of RR Lyrae Stars brighter than 12.5 mag.

J/A+A/442/841/  tablea.dat  (21Kb)

Olsen L.F., Zucca E., Bardelli S., Benoist C., Da Costa L., Jorgensen H.E., Biviano A., Ramella M. Astron. Astrophys., 442, 841-849 (2005) New spectroscopic confirmations of high-redshift galaxy clusters.

J/A+A/442/917/  table1.dat table2.dat table3.dat  (272Kb)

Carraro G., Geisler D., Moitinho A., Baume G., Vazquez R. Astron. Astrophys. 442, 917 (2005) A photometric study of the old open clusters Berkeley 73, Berkeley 75 and Berkeley 25.

J/A+A/442/961/  refs.dat table3.dat table6.dat  (33Kb)

Charbonnel C., Primas F. Astron. Astrophys., 442, 961-992 (2005) The lithium content of the Galactic Halo stars.

J/A+A/442/1023/  stars.dat table3.dat  (17Kb)

Edelmann H., Heber U., Altmann A., Karl C., Lisker T. Astron. Astrophys. 442, 1023 (2005) High resolution spectroscopy of bright subdwarf B stars. I. Radial velocity variables.

J/A+A/442/1135/  notes1.dat notes2.dat refs.dat table1.dat table2.dat  (88Kb)

Liu Q.Z., van Paradijs J., van den Heuvel E.P.J. Astron. Astrophys. 442, 1135 (2005) High-mass X-ray binaries in the Magellanic Clouds.

J/A+A/443/41/  table.dat  (110Kb)

Mengel S., Lehnert M.D., Thatte N., Genzel R. Astron. Astrophys. 443, 41 (2005) Star-formation in NGC 4038/4039 from broad- and narrow band photometry: Cluster Destruction?

J/A+A/443/103/  table1.dat table2.dat  (24Kb)

Carpano S., Wilms J. , Schirmer M., Kendziorra E. Astron. Astrophys. 443, 103 (2005) X-ray properties of NGC 300. I. Global properties of X-ray point sources and their optical counterparts.

J/A+A/443/143/  position.dat refs.dat table1.dat table2.dat  (547Kb)

Groenewegen M.A.T., Blommaert J.A.D.L. Astron. Astrophys. 443, 143 (2005) Mira Variables in the OGLE Bulge fields.

J/A+A/443/157/  table4.dat table5.dat table6.dat table7.dat table8.dat table9.dat  (112Kb)

Paunzen E., Netopil M., Iliev I.K., Maitzen H.M., Claret A., Pintado O.I. Astron. Astrophys., 443, 157-162 (2005) CCD photometric search for peculiar stars in open clusters. VI. NGC 1502, NGC 3105, Stock 16, NGC 6268, NGC 7235 and NGC 7510.

J/A+A/443/413/  table3.dat  (174Kb)

Larsen S.S., Brodie J.P., Strader J. Astron. Astrophys. 443, 413 (2005) Globular clusters in NGC 4365: New K-band imaging and a reassessment of the case for intermediate-age clusters.

J/A+A/443/735 See VI/120/

J/A+A/443/851/  table2.dat table3.dat  (2.0 Mb)

Selman F.J., Melnick J. Astron. Astrophys., 443, 851-861 (2005) The IMF of the field population of 30 Doradus.

J/A+A/444/79/  table8.dat  (20Kb)

Mateos S., Barcons X., Carrera F.J., Ceballos M.T., Hasinger G., Lehmann I., Fabian A.C., Streblyanska A. Astron. Astrophys., 444, 79-99 (2005) XMM-Newton observations of the Lockman Hole. IV: spectra of the brightest AGN.

J/A+A/444/287/  table2.dat table3.dat table4.dat  (12Kb)

Snodgrass C., Fitzsimmons A., Lowry S.C. Astron. Astrophys. 444, 287 (2005) The nuclei of comets 7P/Pons-Winnecke, 14P/Wolf and 92P/Sanguin.

J/A+A/444/305/  table1.dat table2.dat  (287Kb)

Landi E., Bhatia A.K. Astron. Astrophys. 444, 305 (2005) Atomic data and spectral line intensities for Ca XIII.

J/A+A/444/365/  catalog.dat  (20.0 Mb)

White S.D.M., Clowe D.I., Simard L., Rudnick G., De Lucia G., Aragon-Salamanca A., Bender R., Best P., Bremer M., Charlot S., Dalcanton J., Dantel M., Desai V., Fort B., Halliday C., Jablonka P., Kauffmann G., Mellier Y., Milvang-Jense B., Pello R., Poggianti B., Poirier S., Rottgering H., Saglia R., Schneider P., Zaritsky D. Astron. Astrophys. 444, 365 (2005) EDisCS - the ESO Distant Cluster Survey. Sample Definition and Optical Photometry.

J/A+A/444/625/  table1.dat table2.dat table3.dat table4.dat table5.dat table6.dat table7.dat  (26Kb)

Rizvanov N., Nefedjev J. Astron. Astrophys. 444, 625 (2005) Photographic observations of Solar System bodies at the Engelhardt astronomical observatory.

J/A+A/444/643/  table4.dat  (160Kb)

Pourbaix D., Knapp G.R., Szkody P., Ivezic Z., Kleinman S.J., Long D., Snedden S.A., Nitta A., Harvanek M., Krzesinski J., Brewington H.J., Barentine J.C., Neilsen E.H., Brinkman J. Astron. Astrophys. 444, 643 (2005) Candidate spectroscopic binaries in the Sloan Digital Sky Survey

J/A+A/444/685/  table5.dat  (9Kb)

Barrena R., Ramella M., Boschin W., Nonino M., Biviano A., Mediavilla E. Astron. Astrophys. 444, 685 (2005) VGCF detection of galaxy systems at intermediate redshifts.

J/A+A/444/831/  table1.dat table2.dat table3.dat  (15Kb)

Bedin L.R., Piotto G., Baume G., Momany Y., Carraro G., Anderson J., Messineo M., Ortolani S. Astron. Astrophys., 444, 831-836 (2005) Star cluster detection with WFC/ACS in M33. 33 new clusters and 51 candidates.

J/A+A/444/895/  table2.dat table3.dat table4.dat table5.dat  (21Kb)

Steiner J.E., Oliveira, A.S. Astron. Astrophys. 444, 895 (2005) The qWR star HD 45166: I. Observations and system parameters

J/A+A/444/941/  table2.dat  (22Kb)

Paunzen E., Schnell A., Maitzen H.M. Astron. Astrophys., 444, 941-946 (2005) An empirical temperature calibration for the Δa photometric system. I. The B-type stars.

J/A+A/445/69/  table2.dat  (13Kb)

Sohn Y.J., Kang A., Rhee J., Shin M., Chun M.S., Kim H.I. Astron. Astrophys. 445, 69 (2006) Near-IR photometry of asymptotic giant branch stars in the dwarf elliptical galaxy NGC 147.

J/A+A/445/79/  table1.dat table5.dat table6.dat table8.dat table9.dat  (274Kb)

Annibali F., Bressan A., Rampazzo R., Zeilinger W.W. Astron. Astrophys., 445, 79-91 (2006) Nearby early-type galaxies with ionized gas. II. Line-strength indices for 18 additional galaxies.

J/A+A/445/155/  table2.dat table3.dat  (12Kb)

Webb N.A., Wheatley P.J., Barret D. Astron. Astrophys. 445, 155 (2006) XMM-Newton X-ray and optical observations of the globular clusters M 55 and NGC 3201.

J/A+A/445/243/  table2.dat  (21Kb)

Weltevrede P., Edwards R.T., Stappers B.W. Astron. Astrophys. 445, 243 (2006) The subpulse modulation properties of pulsars at 21 cm.

J/A+A/445/361/ table5.dat table6.dat table7.dat table8.dat table9.dat table10.dat table11.dat table12.dat table13.dat table14.dat table15.dat table16.dat (1.5 Mb) Abgrall H., Roueff E.

Astron. Astrophys. 445, 361 (2006) Theoretical calculations of excited rovibrational levels of HD. Term values and transition probabilities of VUV electronic bands.

J/A+A/445/423/  table2.dat table2a.dat table2b.dat  (1.9 Mb)

Fliri J., Riffeser A., Seitz S., Bender R. Astron. Astrophys. 445, 423-439 (2006) The Wendelstein Calar Alto Pixellensing Project (WeCAPP): the M 31 variable star catalogue.

J/A+A/445/441/  table3.dat  (25Kb)

Nieppola E., Tornikoski M., Valtaoja E. Astron. Astrophys., 445, 441-450 (2006) Spectral energy distributions of a large sample of BL Lacertae objects.

J/A+A/445/465/  table1.dat  (21Kb)

Ricci R., Prandoni I., Gruppioni G, Sault R.J., De Zotti G. Astron. Astrophys. 445, 465 (2006) High-frequency observations of the Kuehr sample and epoch-dependent luminosity function of flat-spectrum quasars.

J/A+A/445/493/  table2.dat table3.dat  (704Kb)

Moitinho A., Carraro G., Baume G., Vazquez R.A. Astron. Astrophys. 445, 493 (2006) Open clusters in the Third Galactic Quadrant. II. The intermediate age open clusters NGC 2425 and NGC 2635.

J/A+A/445/503/  table4.dat tablea1.dat  (69Kb)

Reijns R.A., Seitzer P., Arnold R., Freeman K.C., Ingerson T., van den Bosch R.C.E., van de Ven G., de Zeeuw P.T. Astron. Astrophys. 445, 503 (2006) Radial velocities in the globular cluster Omega Centauri.

J/A+A/445/617/  table1.dat table2.dat table3.dat table4.dat table5.dat  (435Kb)

Jurcsik J., Szeidl B., Varadi M., Henden A., Hurta Zs., Lakatos B., Posztobanyi K., Klagyivik P., Sodor A. Astron. Astrophys. 445, 617 (2006) The triple-mode pulsating variable V823 Cas.

J/A+A/445/673/  table4.dat  (9Kb)

Sanz-Forcada J., Favata F., Micela G. Astron. Astrophys. 445, 673 (2006) A compact flare eclipsed in the corona of SV Camelopardalis.

J/A+A/445/765/  table2.dat table3.dat  (898Kb)

Kautsch S.J., Grebel E.K., Barazza F.D., Gallagher J.S.III Astron. Astrophys. 445, 765 (2006) A Catalog of Edge-on Disk Galaxies: From Galaxies with a Bulge to Superthin Galaxies.

J/A+A/445/819/  table1.dat table2.dat  (15Kb)

Ragone C.J., Muriel H., Proust D., Reisenegger A., Quintana H. Astron. Astrophys., 445, 819-825 (2006) Identification and study of systems of galaxies in the Shapley supercluster.

J/A+A/445/857/  table2.dat  (10Kb)

Pompei E., De Carvalho R.R., Iovino A. Astron. Astrophys., 445, 857-867 (2006) The DPOSS II compact group survey: first spectroscopically confirmed candidates.

J/A+A/445/901/  ic1613.dat table3.dat  (45Kb)

Antonello E., Fossati L., Fugazza D., Mantegazza L., Gieren W. Astron. Astrophys. 445, 901 (2006) Variable stars in nearby galaxies. VII. P-L relation in the BVRI bands of Cepheids in IC 1613.

J/A+A/445/939/ notes.dat table1.dat table1.ps table1.tex table2.dat table2.ps table2.tex table3.dat table4.dat (607Kb) Schuster W.J., Moitinho, A., Marquez, A., Parrao, L., Covarrubias, E. Astron. Astrophys. 445, 939 (2006) uvby-beta photometry of high-velocity and metal-poor stars. XI. Ages of halo and old disk stars.

J/A+A/445/971/  table3.dat  (8Kb)

Persi P., Tapia M., Smith H.A. Astron. Astrophys., 445, 971-978 (2006)

Mid-infrared images of the massive star forming region W75 N.

J/A+A/446/19/  table4.dat  (154Kb)

Thomas T., Katgert P. Astron. Astrophys. 446, 19 (2006) The ESO Nearby Abell Cluster Survey. VIII. Morphological and spectral classification of galaxies.

J/A+A/446/97/  table2.dat table3.dat  (20Kb)

Branchesi M., Gioia I.M., Fanti C., Fanti R., Perley R. Astron. Astrophys. 446, 97 (2006) The radio luminosity function of the NEP distant cluster radio galaxies.

J/A+A/446/267/  tableb1.dat tableb2.dat  (15Kb)

Reiners A. Astron. Astrophys. 446, 267 (2006) Rotation- and temperature-dependence of stellar latitudinal differential rotation.

J/A+A/446/485/  table3.dat table4.dat table5.dat  (18Kb)

Guieu S., Dougados C , Monin J-L , Magnier E., Martin E.L Astron. Astrophys. 446, 485 (2006) Seventeen new very low-mass members in Taurus. The brown dwarf deficit revisited.

J/A+A/446/501/  tablea1.dat tablea2.dat tableb.dat  (32Kb)

Franciosini E., Pallavicini R., Sanz-Forcada J. Astron. Astrophys., 446, 501-513 (2006) XMM-Newton observations of the σOrionis cluster. II. Spatial and spectral analysis of the full EPIC field.

J/A+A/446/515/  table1.dat table2.dat  (18Kb)

Phan-Bao N., Bessell M.S. Astron. Astrophys., 446, 515-523 (2006) Spectroscopic distances of nearby ultracool dwarfs.

J/A+A/446/551/  ext7.dat ext15.dat iso2mass.dat  (14.8 Mb)

Jiang B.W., Gao J., Omont A., Schuller F., Simon G. Astron. Astrophys. 446, 551 (2006) Extinctions at 7μm and 15μm from the ISOGAL survey.

J/A+A/446/583/  refs.dat table1.dat  (44Kb)

Libich J., Harmanec P., Vondrak J., Yang S., Hadrava P., Aerts C., De Cat P., Koubsky P., Skoda P., Slechta M., Uytterhoeven K., Mathias P. Astron. Astrophys. 446, 583 (2006) The new orbital elements and properties of epsilon Persei

J/A+A/446/773/  table3.dat table4.dat table5.dat  (24Kb)

Jimenez-Esteban F.M., Garcia-Lario P., Engels D., Perea Calderon J.V. Astron. Astrophys. 446, 773 (2006) An infrared study of galactic OH/IR stars. II. The ”GLMP sample” of red oxygen-rich AGB stars

J/A+A/446/785/  catalog.dat  (570Kb)

Malkov O.Yu., Oblak E., Snegireva E. A., Torra J. Astron. Astrophys. 446, 785 (2006) Catalogue of eclipsing variables.

J/A+A/446/839/  refs.dat table1.dat table2.dat  (64Kb)

Gavazzi G., Boselli A., Cortese L., Arosio I., Gallazzi A., Pedotti P., Carrasco L. Astron. Astrophys. 446, 839 (2006) Hαsurface photometry of galaxies in nearby clusters. V. The survey completion

J/A+A/446/949/  pm/* table1.dat  (9.6 Mb)

Dias W.S., Florio, V., Assafin M., Alessi B.S., Libero V. Astron. Astrophys. 446, 949 (2006) Proper motion determination of open clusters based on the UCAC2 catalogue.

J/A+A/446/1185/  levels.dat tables.dat  (186Kb)

McLaughlin B.M., Hibbert A., Scott M.P., Noble C.J., Burke V.M., Burke P.G. Astron. Astrophys. 446, 1185 (2006) Electron collisions with Fe-peak elements: Fe IV I. Forbidden transitions: 3d(5) - 3d(4)4s and 3d(5) - 3d(4)4p manifolds.

J/A+A/447/71/  tablea1.dat  (32Kb)

Voss R., Gilfanov M. Astron. Astrophys., 447, 71-80 (2006)

The luminosity function of X-ray point sources in Centaurus A.

J/A+A/447/89/  lmc.dat m31.dat m33.dat smc.dat  (33Kb)

Tsalmantza P., Kontizas E., Cambresy L., Genova F., Dapergolas A., Kontizas M. Astron. Astrophys. 447, 89 (2006) Luminous AGB stars in nearby galaxies: A study using Virtual Observatory tools.

J/A+A/447/173/  table8.dat table8.ps table8.tex table9.dat table9.ps table9.tex  (431Kb)

Pauli E.-M., Napiwotzki R., Heber U.,M.altmann, Odenkirchen M. Astron. Astrophys. 447, 173 (2006) 3D kinematics of white dwarfs from the SPY project. II.

J/A+A/447/221/  table1.dat table2.dat  (70Kb)

Beltran M.T., Brand J., Cesaroni R., Fontani F., Pezzuto S., Testi L., Molinari S. Astron. Astrophys., 447, 221-233 (2006) Search for massive protostar candidates in the southern hemisphere. II. Dust continuum emission.

J/A+A/447/245/  table1.dat table1.fit table2.dat table2.fit table3.dat table3.fit  (542Kb)

Wilms J., Nowak M.A., Pottschmidt K., Pooley G.G., Fritz S. Astron. Astrophys. 447, 245 (2006) Long Term Variability of Cyg X-1: IV. Spectral Evolution 1999-2004.

J/A+A/447/473/  table2.dat  (9Kb)

Battinelli P., Demers S. Astron. Astrophys. 447, 473 (2006) The C star population of DDO 190.

J/A+A/447/609/  notes.dat tablea1.dat  (39Kb)

Stanke T., Smith M.D., Gredel R., Khanzadyan T. Astron. Astrophys., 447, 609-622 (2006) An unbiased search for the signatures of protostars in the ρOphiuchi molecular cloud. II. Millimetre continuum observations.

J/A+A/447/685/  abund.dat tab567.dat table4.dat table8.dat tables.tex  (92Kb)

Adelman S.J., Caliskan H., Gulliver A.F., Teker A. Astron. Astrophys. 447, 685 (2006) Elemental abundance analyses with DAO spectrograms XXIX. The Mercury-Manganese stars 53 Tau, beta Tau, gamma Crv, and upsilon Her.

J/A+A/447/735/  dailyns.dat monthns.dat  (588Kb)

Temmer M., Rybak J., Bendik P., Veronig A., Vogler F., Otruba W., Poetzi W., Hanslmeier A. Astron. Astrophys. 447, 735 (2006) Hemispheric Sunspot Numbers Rn and Rs from 1945-2004: Catalogue and N-S asymmetry analysis for solar cycles 18-23.

J/A+A/447/921/  notes.dat table2.dat table3.dat table4a.dat table4b.dat  (69Kb)

Kronberger M., Teutsch P., Alessi B., Steine M., Ferrero L., Graczewski K., Juchert M., Patchick D., Riddle D., Saloranta J., Schoenball M., Watson C. Astron. Astrophys. 447, 921 (2006) New Galactic Open Cluster Candidates from DSS and 2MASS Imagery.

J/A+A/448/L9/  table1.dat  (7Kb)

Moretti A., Perri M., Capalbi M., Angelini L., Hill J.E., Campana S., Burrows D.N., Osborne J.P., Tagliaferri G., Cusumano G., Giommi P., Romano P., Mineo T., Kennea J., Morris D., Nousek J., Pagani C., Racusin J., Abbey A.F., Beardmore A.P., Godet O., Goad M.R., Page K.L., Wells A.A., Chincarini G. Astron. Astrophys. 448, L9 (2006) A refined position catalogue of the Swift XRT afterglows.

J/A+A/448/155/  tablea1.dat tablea2.dat tablea3.dat  (16Kb)

Bergond G., Zepf S.E., Romanowsky A.J., Sharples R.M., Rhode K.L. Astron. Astrophys., 448, 155-164 (2006) Wide-field kinematics of globular clusters in the Leo I group.

J/A+A/448/171/  tablea1.dat tablea2.dat tablea3.dat  (31Kb)

Hilker M. Astron. Astrophys., 448, 171-180 (2006) Probable member stars of the gravitational theory-testing globular clusters AM 1, Pal 3 and Pal 14.

J/A+A/448/181/  apx/* mag1agb.dat mag1pagb.dat mag2agb.dat mag2pagb.dat  (205Kb)

Groenewegen M.A.T. Astron. Astrophys. 448, 181 (2006)

The mid- and far-infrared colours of AGB and post-AGB stars.

J/A+A/448/189/  table3.dat  (88Kb)

de Wit W.J., Bouvier J., Palla F., Cuillandre J.C., James D.J., Kendall T.R., Lodieu N., McCaughrean M.J., Moraux E., Randich S., Testi L. Astron. Astrophys. 448, 189 (2006) Exploring the lower mass function in the young open cluster IC 4665.

J/A+A/448/411/  q.fit q gal.fit u.fit u gal.fit  (13.6 Mb)

Wolleben M., Landecker T.L., Reich W., Wielebinski R. Astron. Astrophys. 448, 411 (2006) An absolutely calibrated survey of polarized emission from the northern sky at 1.4 GHz. Observations and data reduction.

J/A+A/448/579/  pmotions.dat pos 2000.dat table2.dat table3.dat table4.dat  (36Kb)

Carollo D., Bucciarelli B., Hodgkin S.T., Lattanzi M.G., McLean B., Morbidelli R., Smart R.L., Spagna A., Terranegra L. Astron. Astrophys. 448, 579 (2006) The GSC-II-based survey of ancient cool white dwarfs. I. The sample of spectroscopically confirmed WDs.

J/A+A/448/697/  obsgpeg.dat  (27Kb)

Chapellier E., Le Contel D., Le Contel J.-M., Mathias P., Valtier J.-C. Astron. Astrophys. 448, 697 (2006) An hybrid beta Cephei-SPB star in a binary system: gamma Pegasi.

J/A+A/448/779/ efbpss.fe4.user lifetmbp.f table1.dat table2a.dat table2b.dat table2c.dat table2d.dat (21.0 Mb) Nahar S.N. Astron. Astrophys. 448, 779 (2006) Atomic data from the Iron Project. LXI. Radiative E1, E2, E3, and M1 transition probabilities for Fe IV.

J/A+A/448/893/  objects.dat table6.dat table7.dat table8.dat table9.dat  (75Kb)

Lamareille F., Contini T., Le Borgne J.-F., Brinchmann J., Charlot S., Richard J. Astron. Astrophys. 448, 893 (2006) Spectrophotometric properties of galaxies at intermediate redshifts (z 0.2-1.0). I. Sample description, photometric properties and spectral measurements.

J/A+A/448/955/  sdss.dat table1.dat table2.dat  (164Kb)

Izotov Y.I., Stasinska G., Meynet G., Guseva N.G., Thuan T.X. Astron. Astrophys. 448, 955 (2006) The chemical composition of metal-poor emission-line galaxies in the Data Release 3 of the Sloan Digital Sky Survey.

J/A+A/448/983/  table2.dat table3.dat  (37Kb)

Rejkuba M., Da Costa G.S., Jerjen, H., Zoccali, M., Binggeli B. Astron. Astrophys. 448, 983 (2006) Dwarf elliptical galaxies in Centaurus A group: stellar populations in AM 1339-445 and AM1343-452.

J/A+A/448/1001/  table1.dat  (10Kb)

Jilinski E., Daflon S., Cunba K., de la Raze R. Astron. Astrophys. 448, 1001 (2006) Radial velocity measurements of B stars in the Scorpius-Centaurus association.

J/A+A/448/1075/  table1.dat  (17Kb)

Shakhovskoj D., Grinin V., Rostopchina, A., Schertl D., Hofmann K.-H., Weigelt G., Balega Yu., Kiyaeva O., Melnikov S. Astron. Astrophys. 448, 1075 (2006) Photopolarimetric activity and circumstellar environment of the young binary system DF Tau.

J/A+A/448/1107/  table1.dat  (141Kb)

Dobrotka A., Friedjung M., Retter A., Hric L., Novak R. Astron. Astrophys. 448, 1107 (2006) The possible orbital period of the nova V1493 Aquilae.

J/A+A/448/1235 See I/300/

J/A+A/448/1247/  table4.dat table5.dat  (1.1 Mb)

Misanovic Z., Pietsch W., Haberl F., Ehle M., Hatzidimitriou D., Trinchieri G Astron. Astrophys. 448, 1247 (2006) XMM-Newton survey of the Local Group galaxy M33. Variability of the detected sources.

J/A+A/449/127/  stars.dat table1.dat table3.dat  (253Kb)

Zhang H.W., Zhao G. Astron. Astrophys. 449, 127 (2006) Chemical abundances of 32 mildly metal-poor stars.

J/A+A/449/143/  sources.dat table1.dat table2.dat  (24Kb)

Ma J., Zhou X., Burstein D., Yang Y., Fan Z., Chen J., Jiang Z., Wu Z., Wu J., Zhang T. Astron. Astrophys. 449, 143 (2006) Age estimations of M31 globular clusters from their spectral energy distributions.

J/A+A/449/305/  table1.dat  (27Kb)

Aerts C., De Cat P., De Ridder J., Van Winckel H., Raskin G., Davignon G., Uytterhoeven K. Astron. Astrophys. 449, 305 (2006) Multiperiodicity in the large-amplitude rapidly-rotating beta Cephei star HD 203664.

J/A+A/449/425/  table1.dat  (685Kb)

Mickaelian A.M., Hovhannisyan L.R., Engels D., Hagen H.-J., Voges W. Astron. Astrophys. 449, 425 (2006) Optical identification of ROSAT-FSC sources.

J/A+A/449/435 See I/303/

J/A+A/449/461/  table1.dat  (12Kb)

Boschin W., Girardi M., Spolaor M., Barrena R. Astron. Astrophys. 449, 461 (2006) Internal dynamics of the radio-halo cluster A2744.

J/A+A/449/583/  jfay.dat jfor.dat jforfir.dat joli.dat  (392Kb)

Koike C., Mutschke H., Suto H., Naoi T., Chihara H., Henning T., Jager C., Tsuchiyama A., Dorschner J., Okuda H. Astron. Astrophys. 449, 583 (2006) Temperature effects on the mid-and far-infrared spectra of olivine particles.

J/A+A/449/723/  compar.dat planets.dat stars.dat  (565Kb)

Gilli G., Israelian G., Ecuvillon A., Santos N.C., Mayor M. Astron. Astrophys. 449, 723 (2006) Abundances of refractory elements in the atmospheres of stars with extrasolar planets.

J/A+A/449/855/  table8.dat table9.dat table10.dat table11.dat  (1.0 Mb)

Coudert L.H., Roueff E. Astron. Astrophys. 449, 855 (2006) Linelists for NH3, NH2D, ND2H, and ND3 with quadrupole coupling hyperfine components.

J/A+A/449/929/  refs.dat table4.dat  (24Kb)

Gavazzi G., O’neil K., Boselli A., Van Driel W. Astron. Astrophys., 449, 929-935 (2006) HI observations of galaxies. II. The Coma Supercluster.

J/A+A/449/937/  refs.dat table1.dat table3.dat table4.dat table5.dat  (42Kb)

Sulentic J.W., Verdes-Montenegro L., Bergond G., Lisenfeld U., Durbala A., Espada D., Garcia E., Leon S., Sabater J., Verley S., Casanova V., Sota A. Astron. Astrophys. 449, 937 (2006) The AMIGA sample of isolated galaxies. II. Morphological refinement.

J/A+A/449/951/  table5.dat table6.dat table7.dat  (5.6 Mb)

Grazian A., Fontana A., De Santis C., Nonino M., Salimbeni S., Giallongo E., Cristiani S., Gallozzi S., Vanzella E. Astron. Astrophys. 449, 951 (2006) The GOODS-MUSIC sample: a multicolour catalog of near-IR selected galaxies in the GOODS-South field.

J/A+A/450/39/  table2.dat  (121Kb)

Gu M.F., Lee C.-U., Pak Soojong, Yim H.S., Fletcher A.B. Astron. Astrophys. 450, 39 (2006) Multi-colour optical monitoring of eight red blazars.

J/A+A/450/105/  table1.dat  (24Kb)

Bica E., Bonatto C., Barbuy B., Ortolani S. Astron. Astrophys. 450, 105 (2006) Globular cluster system and Milky Way properties revisited.

J/A+A/450/147/  table4.dat  (54Kb)

Lodieu N., Bouvier J., James D.J., de Wit W.-J., Palla F., McCaughrean M.J., Cuillandre J.-C. Astron. Astrophys. 450, 147 (2006)

A deep wide-field optical survey in the young open cluster Collinder 359.

J/A+A/450/253/  table5.dat  (26Kb)

Maercker M., Burton M.G., Wright C.M. Astron. Astrophys. 450, 253 (2006) L-band (3.5 micron) IR-excess in massive star formation. II. RCW 57/NGC 3576.

J/A+A/450/681/  notes.dat table2.dat table5.dat  (52Kb)

Tokovinin A., Thomas S., Sterzik M., Udry S. Astron. Astrophys. 450, 681 (2006) Tertiary companions to close spectroscopic binaries.

J/A+A/450/715/  table1.dat  (72Kb)

Rodriguez E., Amado P.J., Suarez J.C., Moya A., Dupret M.A., Poretti E., Grigahcene A., Costa V., Lopez-Gonzalez M.J. Astron. Astrophys. 450, 715 (2006) The frequency ratio method and the new multiperiodic Gamma Dor star HD 218427.

J/A+A/450/735/  table2.dat table3.dat  (880Kb)

Masana E., Jordi C., Ribas I. Astron. Astrophys. 450, 735 (2006) Effective temperature scale and bolometric corrections from 2MASS photometry.

J/A+A/450/763/  table1.dat table2.dat  (33Kb)

Kochukhov O., Bagnulo S. Astron. Astrophys., 450, 763-775 (2006) Evolutionary state of magnetic chemically peculiar stars.

J/A+A/450/855/ eep03.dat sp03.dat sxy2dp03.dat sxy2p03.dat xciop03.dat yciop03.dat zciop03.dat (226Kb) Capitaine N., Wallace P.T. Astron. Astrophys. 450, 855 (2006) High precision methods for locating the celestial intermediate pole and origin.

J/A+A/450/873/  table1.dat table2.dat table3.dat tables.tex  (222Kb)

Mennickent R.E., Gieren W., Soszynski I., Pietrzynski G. Astron. Astrophys. 450, 873 (2006) The Araucaria project. Bright variable stars in NGC 6822 from a wide-field imaging survey.

J/A+A/450/959/  table2.dat table3.dat table4.dat  (17Kb)

Orienti M., Dallacasa D., Tinti S., Stanghellini C. Astron. Astrophys. 450, 959 (2006) VLBA images of High Frequency Peakers.

J/A+A/450/993/  tablea1.dat tableb1.dat tablec1.dat  (78Kb)

Pillitteri I., Micela G., Damiani F., Sciortino S. Astron. Astrophys., 450, 993-1004 (2006) Deep X-ray survey of the young open cluster NGC 2516 with XMM-Newton.

J/A+A/450/1149/  table8.dat  (67Kb)

Telting J.H., Ostensen R.H. Astron. Astrophys. 450, 1149 (2006) Time-resolved spectroscopy of the bright sdBV Balloon 090100001.

J/A+A/450/1249/  table1.dat table2.dat table3.dat table4.dat table5.dat tables.ps tables.tex  (751Kb)

Aggarwal K.M., Keenan F.P. Astron. Astrophys. 450, 1249 (2006) Electron impact excitation of Fe XVI: radiative and excitation rates.

J/A+A/451/57/  table3.dat  (11Kb)

Marcillac D., Elbaz D., Chary R.R., Dickinson M., Galliano F., Morrison G. Astron. Astrophys., 451, 57-69 (2006) Mid infrared properties of distant infrared luminous galaxies.

J/A+A/451/431/  table2.dat table3.dat  (16Kb)

Livanou E., Kontizas M., Gonidakis I., Kontizas E., Maragoudaki F., Oliver S., Efstathiou A., Klein U. Astron. Astrophys. 451, 431 (2006) Star-burst regions in the LMC. Mapping and multi-wavelength properties of stellar complexes.

J/A+A/451/435/  table2.dat tablea1.dat  (452Kb)

Montagni F., Maselli A., Massaro E., Nesci R., Sclavi S., Maesano M. Astron. Astrophys., 451, 435-442 (2006)

The intra-night optical variability of the bright BL Lac object S5 0716+714.

J/A+A/451/457/  table1.dat  (37Kb)

Tozzi P., Gilli R., Mainieri V., Norman C., Risaliti G., Rosati P., Bergeron J., Borgani S., Giacconi R., Hasinger G., Nonino M., Streblyanska A., Szokoly G., Wang J.X., Zheng W. Astron. Astrophys., 451, 457-474 (2006) X-ray spectral properties of active galactic nuclei in the Chandra Deep Field South.

J/A+A/451/835/  notes.dat table1.dat table2.dat  (17Kb)

Hatzidimitriou D., Pietsch W., Misanovic Z., Reig P., Haberl F. Astron. Astrophys. 451, 835 (2006) Spectroscopy of the brightest optical counterparts of X-ray sources in the direction of M 31 and M 33.

J/A+A/451/901/  table2.dat table3a.dat table3b.dat table5.dat  (398Kb)

Frolov V.N., Ananjevskaja J.K., Jilinski E.G., Gorshanov D.L., Bronnikova N.M. Astron. Astrophys. 451, 901 (2006) Proper motions and CCD-photometry of stars in the region of the open cluster Trumpler 2.

J/A+A/451/1053/  sp/* table1.dat table4.dat table5.dat table7a.dat table7b.dat table7c.dat  (35.0 Mb)

Fremat Y., Neiner C., Hubert A.-M., Floquet M., Zorec J., Janot-Pacheco E., Renan de Medeiros J. Astron. Astrophys. 451, 1053 (2006) Fundamental parameters of Be stars located in the seismology fields of COROT.

J/A+A/451/1125/  levels.dat tables.dat  (158Kb)

Lique F., Spielfiedel A., Cernicharo J. Astron. Astrophys. 451, 1125 (2006) Rotational excitation of carbon monosulfite by collisions with He.

J/A+A/451/1159/  catalog.dat  (19.2 Mb)

Adami C., Picat J.P., Savine C., Mazure A., West M.J., Cuillandre J.-C., Pello R., Biviano A., Conselice C.J., Durret F., Gallagher J.S.III, Gregg M., Moreau C., Ulmer M. Astron. Astrophys., 451, 1159-1170 (2006) Deep and wide field imaging of the Coma cluster: the data.

J/A+A/452/119/  cat/* catalog.dat desccat.dat fields.dat ima/*  (199.4 Mb)

Olsen L.F., Miralles J.-M., da Costa L., Benoist C., Vandame B., Rengelink R., Rite C., Scodeggio M., Slijkhuis R., Wicenec A., Zaggia S. Astron. Astrophys. 452, 119 (2006) ESO imaging survey: infrared observations of CDF-S and HDF-S.

J/A+A/452/179/  table2.dat  (40Kb)

Chiosi E., Vallenari A., Held E.V., Rizzi L., Moretti A. Astron. Astrophys. 452, 179 (2006) Age distribution of young clusters and field stars in the SMC.

J/A+A/452/245/  tablec1.dat tablec2.dat tablec3.dat  (27Kb)

Natta A., Testi L., Randich S. Astron. Astrophys., 452, 245-252 (2006) Accretion in the ρ-Ophiuchi pre-main sequence stars.

J/A+A/452/503/  table1.dat  (10Kb)

Kotnik-Karuza D., Friedjung M., Whitelock P.A., Marang F., Exter K., Keenan F.P., Pollacco D.L. Astron. Astrophys. 452, 503 (2006) The effect of dust obscuration in RR Tel on optical and IR long-term photometry and Fe II emission lines.

J/A+A/452/567/  fluxlr.dat halpha.dat oi.dat  (94Kb)

Munari U., Siviero A., Navasardyan H., Dallaporta S. Astron. Astrophys. 452, 567 (2006) BVI photometry and the spectroscopy of Nova Scuti 2005 N.2.

J/A+A/452/739/  table3.dat  (11Kb)

Sanchez-Sutil J.R., Munoz-Arjonilla A.J., Marti J., Garrido J.L., Perez-Ramirez D., Luque-Escamilla P. Astron. Astrophys., 452, 739-742 (2006) A catalogue of ultra-luminous X-ray source coincidences with FIRST radio sources

J/A+A/452/959/  table10.dat table11.dat  (37Kb)

Kallrath J., Milone E.F., Breinhorst R.A., Wilson R.E., Schnell A., Purgathofer A. Astron. Astrophys. 452, 959 (2006) V781 Tauri: A W Ursae Majoris Binary with Decreasing Period.

J/A+A/452/975/  table1.dat  (501Kb)

Axelsson M., Borgonovo L., Larsson S.

Astron. Astrophys. 452, 975 (2006) Probing the temporal variability of Cygnus X-1 into the soft state.

J/A+A/452/1113/  table5.dat table7.dat  (70Kb)

Hudson C.E., Bell K.L. Astron. Astrophys. 452, 1113 (2006) Fine structure effective collision strengths for the electron impact excitation of S V.

J/A+A/453/101/  table6.dat table7.dat table8.dat table9.dat  (2.3 Mb)

Lyra W., Moitinho A., van der Bliek N.S., Alves J.F. Astron. Astrophys. 453, 101 (2006) On the difference between nuclear and contraction ages.

J/A+A/453/155/  table4.dat  (15Kb)

Merand A., Kervella P., Coude du Foresto V., Perrin G., Ridgway S.T., Aufdenberg J.P., ten Brummelaar T.A., McAlister H.A., Sturmann L., Sturmann2 J., Turner N.H., Berger D.H. Astron. Astrophys. 453, 155 (2006) Extended envelopes around Galactic Cepheids. II. Polaris and delta Cep from near-infrared interferometry with CHARA/FLUOR

J/A+A/453/309/  table1.dat table345.dat  (20Kb)

Nardetto N., Mourard D., Kervella P., Mathias P., Merand A., Bersier D. Astron. Astrophys., 453, 309-319 (2006) High resolution spectroscopy for Cepheids distance determination. I. Line asymmetry.

J/A+A/453/349/  table2.dat  (14Kb)

Veiga C.H. Astron. Astrophys. 453, 349 (2006) CCD positions for eight Jovian irregular satellites.

J/A+A/453/371 See II/270/

J/A+A/453/493/  tablec1.dat tablec2.dat  (20Kb)

Serra P., Trager S.C., Van der Hulst J.M., Oosterloo T.A., Morganti R. Astron. Astrophys. 453, 493 (2006) IC 4200: a gas-rich early-type galaxy formed via a major merger.

J/A+A/453/517/  table3.dat  (15Kb)

Testor G., Lemaire J.L., Field D., Diana S. Astron. Astrophys. 453, 517 (2006) VLT/NACO near-infrared imaging and spectroscopy of N159A in the LMC HII complex N159.

J/A+A/453/535/  tablea1.dat  (15Kb)

Schuller F., Omont A., Glass I.S., Schultheis M., Egan M.P., Price S.D. Astron. Astrophys. 453, 535 (2006) Recent star formation in the inner Galactic Bulge seen by ISOGAL. I. Classification of bright mid-IR sources in a test field.

J/A+A/453/635/  table1.dat  (52.1 Mb)

Marshall D.J., Robin A.C., Reyle C., Schultheis M., Picaud S. Astron. Astrophys. 453, 635 (2006) Modelling the Galactic Interstellar Extinction Distribution in Three Dimensions.

J/A+A/453/769/  all/* grid.dat z007.dat z010.dat  (125.3 Mb)

Claret A. Astron. Astrophys. 453, 769 (2006) New grids of stellar models including tidal-evolution constants up to carbon burning. III. From 0.8 to 125M⊙: the Large Magellanic Cloud (Z=0.007-0.01).

J/A+A/453/911/  table2.dat  (8Kb)

Felli M., Massi F., Robberto M., Cesaroni R. Astron. Astrophys., 453, 911-922 (2006) New signposts of massive star formation in the S235A-B region.

J/A+A/453/923/  notes.dat tablea1.dat tablea2.dat  (45Kb)

Kiss Z.T., Toth L.V., Krause O., Kun M, Stickel M. Astron. Astrophys. 453, 923 (2006) Star formation in the Cepheus Flare region - Implications from morphology and infrared properties of optically selected clouds.

J/A+A/453/1003/  refs.dat table1.dat table2.dat  (35Kb)

Thompson M.A., Hatchell J., Walsh A.J., Macdonald G.H., Millar T.J.

Astron. Astrophys., 453, 1003-1026 (2006) A SCUBA imaging survey of ultracompact HII regions. The environments of massive star formation.

J/A+A/454/171/ table5.dat table6.dat table7.dat table8.dat table9.dat table10.dat table11.dat table12.dat (145Kb) Paunzen E., Netopil M., Iliev I.K., Maitzen H.M., Claret A., Pintado O.I. Astron. Astrophys., 454, 171-178 (2006) CCD photometric search for peculiar stars in open clusters. VII. Berkeley 11, Berkeley 94, Haffner 15, Lynga 1, NGC 6031, NGC 6405, NGC 6834 and Ruprecht 130.

J/A+A/454/179/  table1.dat table2.dat table3.dat  (29Kb)

Netopil M., Maitzen H.M., Paunzen E., Claret A. Astron. Astrophys., 454, 179-184 (2006) Photometric survey of marginally investigated open clusters. I. Basel 11b, King 14, Czernik 43.

J/A+A/454/185/  table3.dat  (14Kb)

Hamadache C., Le Guillou L., Tisserand P., Afonso C., Albert J.N., Andersen J., Ansari R., Aubourg E., Bareyre P., Beaulieu J.P., Charlot X., Coutures C., Ferlet R., Fouque P., Glicenstein J.F., Goldman B., Gould A., Graff D., Gros M., Haissinski J., De Kat J., Lesquoy E., Loup C., Magneville C., Marquette J.B., Maurice E., Maury A., Milsztajn A., Moniez M., Palanque-Delabrouille N., Perdereau O., Rahal Y.R., Rich J., Spiro M., Vidal-Madjar A., Vigroux L., Zylberajch S. Astron. Astrophys., 454, 185-199 (2006) Galactic Bulge microlensing optical depth from EROS-2.

J/A+A/454/257/  table5.dat table7.dat  (89Kb)

Nagy A., Kovacs G. Astron. Astrophys., 454, 257-264 (2006) On the incidence rate of first overtone Blazhko stars in the Large Magellanic Cloud.

J/A+A/454/277/  table2.dat table3.dat  (226Kb)

Piirola V., Berdyugin A., Coyne G.V., Efimov Yu.S., Shakhovskoy N.M. Astron. Astrophys. 454, 277 (2006) UBVRI polarimetry of the massive interacting binary SX Cas: modeling the electron-scattering circumstellar envelope.

J/A+A/454/295/  table1.dat  (55Kb)

Froehlich H.-E., Kroll P., Strassmeier K.G. Astron. Astrophys. 454, 295 (2006) The RS CVn binary HK Lacertae: long-term photometry from Sonneberg Sky-Patrol plates.

J/A+A/454/301/ table2.dat table3.dat table4.dat table5.dat table6.dat table7.dat tables.ps tables.tex (77Kb) Frasca A., Guillout P., Marilli E., Freire Ferrero R., Biazzo K., Klutsch A. Astron. Astrophys. 454, 301 (2006) Newly discovered active binaries in the RasTyc sample of stellar X-ray sources. I. Orbital and physical parameters of six new binaries.

J/A+A/454/333/  table1.dat  (591Kb)

Castelli F., Kurucz R.L. Astron. Astrophys. 454, 333 (2006) Computed Hβindices from ATLAS9 model atmospheres.

J/A+A/454/379/  phoebe.dat  (10Kb)

Qiao R.C., Tang Z.H., Shen K.X., Dourneau G., Yan Y.R., Yu Y., Wang S.H., Liu J.R. Astron. Astrophys. 454, 379 (2006) CCD astrometric observations of Phoebe in 2003-2004.

J/A+A/454/423/  table2.dat table2.fit  (279Kb)

Vanzella E., Cristiani S., Dickinson M., Kuntschner H., Nonino M., Rettura A., Rosati P., Vernet J., Cesarsky C., Ferguson H.C., Fosbury R.A.E., Giavalisco A., Grazian M., Haase J., Moustakas L.A., Popesso P., Renzini A., Stern D., GOODS Team Astron. Astrophys. 454, 423 (2006) The Great Observatories Origins Deep Survey VLT/FORS2 Spectroscopy in the GOODS-South Field: part II.

J/A+A/454/453/  tablec1.dat  (16Kb)

Della Valle A., Mazzei P., Bettoni D., Aussel H., De Zotti G. Astron. Astrophys., 454, 453-458 (2006) A multi-wavelength study of the IRAS Deep Survey galaxy sample. I. Optical observations. The redshift distribution.

J/A+A/454/581/ (3Kb) Desidera S., Gratton R.G., Lucatello S., Claudi R.U. Astron. Astrophys., 454, 581-593 (2006) Abundance difference between components of wide binaries. II. The southern sample.

J/A+A/454/595/  table6.dat  (11Kb)

Gondoin P. Astron. Astrophys. 454, 595 (2006) X-ray emission from T Tauri stars in the Lupus 3 star-forming region.

J/A+A/454/683/  table1.dat  (24Kb)

Veiga C.H., Bourget P. Astron. Astrophys. 454, 683 (2006) Puck satellite Earth-based observations.

J/A+A/454/717/  table2.dat table3.dat  (71Kb)

Kang A., Sohn Y.-J., Kim H.-I., Rhee J., Kim J.-W., Hwang N., Lee M.G, Kim Y.-C., Chun M.-S. Astron. Astrophys. 464, 717 (2006) The evolved asymptotic giant branch stars in the central bar of the dwarf irregular galaxy NGC 6822.

J/A+A/454/855/  table1.dat  (32Kb)

Pavlovski, K., Burki, G., Mimica P. Astron. Astrophys. 454, 855 (2006) Indirect imaging of an aacretion disk rim in the long-period interacting binary W Crucis.

J/A+A/454/895/  stars.dat table5.dat table15.dat  (672Kb)

Allen D.M., Barbuy B. Astron. Astrophys. 464, 895 (2006) Analysis of 26 barium stars I. Abundances.

J/A+A/454/917/  stars.dat table2.dat table6.dat  (289Kb)

Allen D.M., Barbuy B. Astron. Astrophys. 464, 917 (2006) Analysis of 26 barium stars. II. Contributions of s-, r-, and p-processes in the production of heavy elements

J/A+A/454/1047/  table3.dat table6.dat  (408Kb)

Sana H., Gosset E., Rauw G., Sung H., Vreux J.-M. Astron. Astrophys. 454, 1047 (2006) An XMM-Newton view of the NGC 6231 young open cluster: I. The catalogue.

J/A+A/455/L1/  table3.dat table4.dat table5.dat table6.dat table7.dat table8.dat  (38Kb)

Paraficz D., Hjorth J., Burud I., Jakobsson P., Eliasdottir A. Astron. Astrophys. 455, L1 (2006) Microlensing variability in time-delay quasars.

J/A+A/455/45/  table1.dat  (9Kb)

Girardi M., Boschin W., Barrena R. Astron. Astrophys., 455, 45-59 (2006) Internal dynamics of the massive cluster Abell 697: a multiwavelength analysis.

J/A+A/455/107/  table2.dat  (10Kb)

Flores H., Hammer F., Puech M., Amram P., Balkowski C. Astron. Astrophys., 455, 107-118 (2006) 3D spectroscopy with VLT/GIRAFFE. I. The true Tully-Fisher relationship at z 0.6.

J/A+A/455/247/  table2.dat  (116Kb)

Tian B., Deng L., Han Z., Zhang X.B. Astron. Astrophys. 455, 247 (2006) The blue stragglers formed via mass transfer in old open clusters.

J/A+A/455/259/  table1.dat  (17Kb)

Ausseloos M., Aerts C., Lefever K., Davis J., Harmanec P. Astron. Astrophys. 455, 259 (2006) High-precision elements of double-lined spectroscopic binaries from combined interferometry and spectroscopy. Application to the beta Cephei star beta Centauri.

J/A+A/455/271/  table2.dat table3.dat table4.dat  (91Kb)

Gratton R.G., Lucatello S., Bragaglia A., Carretta E., Momany Y., Pancino E., Valenti E. Astron. Astrophys. 455, 271 (2006) Na-O anticorrelation and HB. III. The abundances of NGC 6441 from FLAMES-UVES spectra.

J/A+A/455/303/  table1.dat  (28Kb)

Joshi S., Mary D.L., Martinez P., Kurtz D.W., Girish V., Seetha S., Sagar R., Ashoka B.N. Astron. Astrophys., 455, 303-313 (2006) The Nainital-Cape Survey. II. Report for pulsation in five chemically peculiar A-type stars and presentation of 140 null results.

J/A+A/455/453/  table3.dat table4.dat  (64Kb)

Chies-Santos A.L., Pastoriza M.G., Santiago B.X., Forbes D.A. Astron. Astrophys. 455, 453 (2006) The globular cluster system of NGC 5846 revisited: colours, sizes and X-ray counterparts.

J/A+A/455/765 See III/237A/

J/A+A/455/773 See VII/248/

J/A+A/455/813/  refs.dat table1.dat table2.dat  (18Kb)

De Pasquale M., Piro L., Gendre B., Amati L., Antonelli L.A., Costa E., Feroci M., Frontera F., Nicastro L., Soffitta P., in’t Zand J. Astron. Astrophys. 455, 813 (2006) The BeppoSAX catalog of GRB X-ray afterglow observations.

J/A+A/455/903/  table1.dat table3.dat table4.dat table6.dat  (240Kb)

Flaccomio E., Micela G., Sciortino S. Astron. Astrophys., 455, 903-921 (2006) ACIS-I observations of NGC 2264. Membership and X-ray properties of PMS stars.

J/A+A/455/971/  table1.dat table2.dat table7.dat table7.tex  (148Kb)

Requena-Torres M.A., Martin-Pintado J., Rodriguez-Franco A., Martin S., Rodriguez-Fernandez N.J., de Vicente P. Astron. Astrophys. 455, 971 (2006) Organic molecules in the Galactic center.

J/A+A/455/1037/  table3.dat  (125Kb)

Linnell A.P., Harmanec P., Koubsky P., Bozic H., Yang S., Ruzdjak D., Sudar D., Libich J., Eenens P., Krpata J., Wolf M., Skoda P., Slechta M. Astron. Astrophys. 455, 1037 (2006) Properties and nature of Be stars. 24. Better data and model for the Be+F binary V360 Lacertae

J/A+A/455/1157/  table1.dat table6.dat table7.dat table8.dat  (1.2 Mb)

Jonauskas V., Bogdanovich P., Keenan F.P., Kisielius R., Foord M.E., Heeter R.F., Rose S.J., Ferland G.J., Norrington P.H. Astron. Astrophys. 455, 1157 (2006) Energy levels and transition probabilities for boron-like Fe XXII.

J/A+A/455/1161/  table3.dat  (10Kb)

Bizzocchi L., Degli Esposti C., Dore L. Astron. Astrophys., 455, 1161-1164 (2006) Improved rest frequencies for the submillimetre-wave spectrum of SiN.

J/A+A/455/1165/  notes.dat refs.dat table1.dat table1.tex  (252Kb)

Liu Q.Z., van Paradijs J., van den Heuvel E.P.J. Astron. Astrophys. 455, 1165 (2006) Catalogue of high-mass X-ray binaries in the Galaxy (4th edition).

J/A+A/456/261/  table2.dat table3.dat table4.dat  (56Kb)

Rodriguez E., Costa V., Zhou A.-Y., Grigahcene A., Dupret M.A., Suarez J.C., Moya A., Lopez-Gonzalez M.J., Wei J.-Y., Fan Y. Astron. Astrophys. 456, 261 (2006) Asteroseismology of the new multiperiodic Gamma Dor variable HD 239276.

J/A+A/456/451/ table3.dat table3a.dat table4.dat table4a.dat table5.dat table6.dat tablea1.dat tablea1a.dat tablea2.dat tablea2a.dat (794Kb) Leisy P., Dennefeld M. Astron. Astrophys., 456, 451-466 (2006) Planetary nebulae in the Magellanic Clouds. II. Abundances and element production.

J/A+A/456/467/  table1.dat  (19Kb)

Focardi P., Zitelli V., Marinoni S., Kelm B. Astron. Astrophys. 456, 467 (2006) A new sample of bright galaxy pairs in UZC.

J/A+A/456/599/  figa2.dat figa3.dat figa4.dat figa5.dat figa6.dat table1.dat tablea1.dat  (304Kb)

Papadaki C., Boffin H.M.J., Sterken C., Stanishev V., Cuypers J., Boumis P., Akras S., Alikakos J.

Astron. Astrophys. 456, 599 (2006) Photometric study of selected cataclysmic variables.

J/A+A/456/623/  notes.dat table4.dat table5.dat table6.dat table7.dat table8.dat  (85Kb)

Evans C.J., Lennon D.J., Smartt S.J., Trundle C. Astron. Astrophys., 456, 623-638 (2006) The VLT-FLAMES survey of massive stars: observations centered on the Magellanic Cloud clusters NGC 330, NGC 346, NGC 2004, and the N11 region.

J/A+A/456/651/  table1.dat  (1.4 Mb)

Bruntt H., Southworth J., Torres G., Penny A.J., Clausen J.V., Buzasi D.L. Astron. Astrophys., 456, 651-658 (2006) Eclipsing binaries observed with the WIRE satellite. I. Discovery and photometric analysis of the new bright A0IV eclipsing binary ψCentauri.

J/A+A/456/791/  tablea1.dat tablea2.dat  (241Kb)

Tasse C., Cohen A.S., Roettgering H.J.A., Kassim N.E., Pierre M., Perley R., Best P., Birkinshaw M., Bremer M., Liang H. Astron. Astrophys., 456, 791-800 (2006) New results from the low-frequency counterpart of the XMM large scale structure survey.

J/A+A/456/881/  cat/* deep1.dat deep2.dat deep3.dat desccat.dat fields.dat ima/*  (28.1 Mb)

Olsen L.F., Miralles J.-M., da Costa L., Madejsky R., Jorgensen H.E., Mignano A., Arnouts S., Benoist C., Dietrich J.P., Slijkhuis R., Zaggia S. Astron. Astrophys. 456, 881 (2006) ESO imaging survey: infrared deep public survey.

J/A+A/456/905/  table2.dat table3.dat  (25Kb)

Demers S., Battinelli P., Artigau E. Astron. Astrophys. 456, 905 (2006) C stars in the outer spheroid of NGC 6822.

J/A+A/456/953/  table1.dat  (16Kb)

Bennert N., Jungwiert B., Komossa S., Haas M., Chini R. Astron. Astrophys. 456, 953 (2006) Size and properties of the narrow-line region in Seyfert-2 galaxies from spatially-resolved optical spectroscopy

J/A+A/456/977/  table1.dat table2.dat  (42Kb)

Marino A., Micela G., Pillitteri I., Peres G. Astron. Astrophys., 456, 977-983 (2006) X-ray variability of NGC 2516 stars in the XMM-Newton observations.

J/A+A/456/985/  table5.dat  (43Kb)

Galleti S., Federici L., Bellazzini M., Buzzoni A., Fusi Pecci F. Astron. Astrophys. 456, 985 (2006) An updated survey of globular clusters in M 31. I. Classification and radial velocity for 76 candidate clusters.

J/A+A/457/79/  tablea1.dat tablea2.dat  (23Kb)

Gavignaud I., Bongiorno A., Paltani S., Mathez G., Zamorani G., Moller P., Picat J.P., Le Brun V., Marano B., Le Fevre O., Bottini D., Garilli B., Maccagni D., Scaramella R., Scodeggio M., Tresse L., Vettolani G., Zanichelli A., Adami C., Arnaboldi M., Arnouts S., Bardelli S., Bolzonella M., Cappi A., Charlot S., Ciliegi P., Contini T., Foucaud S., Franzetti P., Guzzo L., Ilbert O., Iovino A., Mccracken H.J., Marinoni C., Mazure A., Meneux B., Merighi R., Pello R., Pollo A., Pozzetti L., Radovich M., Zucca E., Bondi M., Busarello G., Cucciati O., De La Torre S., Gregorini L., Lamareille F., Mellier Y., Merluzzi P., Ripepi V., Rizzo D., Vergani D. Astron. Astrophys., 457, 79-90 (2006) The VIMOS VLT Deep Survey: the faint type-1 AGN sample.

J/A+A/457/265/  table7.dat table8.dat table9.dat  (28Kb)

Dufton P.L., Smartt S.J., Lee J.K, Ryans R.S.I., Hunter I., Evans C.J., Herrero A., Trundle C., Lennon D.J., Irwin M.J., Kaufer A. Astron. Astrophys. 457, 265 (2006) The VLT-FLAMES Survey of Massive Stars: Stellar parameters and rotational velocities in NGC 3293, NGC 4755 and NGC 6611.

J/A+A/457/517/  table4.dat table6.dat  (34Kb)

Prandoni I., Parma P., Wieringa M.H., De Ruiter H.R., Gregorini L., Mignano A., Vettolani G., Ekers R.D. Astron. Astrophys., 457, 517-529 (2006) The ATESP 5 GHz radio survey. I. Source counts and spectral index properties of the faint radio population.

J/A+A/457/637/  table2.dat table3.dat  (13Kb)

Martins F., Plez B. Astron. Astrophys. 457, 637 (2006) UBVJHK synthetic photometry of Galactic O stars.

J/A+A/457/651/  tablea1.dat  (22Kb)

Morel T., Butler K., Aerts C., Neiner C., Briquet M. Astron. Astrophys. 457, 651 (2006) Abundance analysis of prime B-type targets for asteroseismology I. Nitrogen excess in slowly-rotating beta Cephei stars.

J/A+A/457/721/ f.allowed.fe18.user f.forbid.fe18.user lifetime.f set.dat table3a.dat table3b.dat table4a.dat table4b.dat table7.dat table8a.dat table8b.dat (25.5 Mb) Nahar S.N. Astron. Astrophys. 457, 721 (2006) Atomic data from the Iron Project. LXII. Allowed and forbidden transitions in Fe XVIII in relativistic Breit-Pauli approximation.

J/A+A/457/1033/  table.dat  (22Kb)

Lehmann H., Tsymbal V., Mkrtichian D.E., Fraga L. Astron. Astrophys. 457, 1033 (2006) The helium-weak silicon star HR 7224 I. Radial velocity and line profile variations.

J/A+A/458/21/  table1.dat  (9Kb)

Falanga M., Goetz D., Goldoni P., Farinelli R., Goldwurm A., Mereghetti S., Bazzano A., Stella L. Astron. Astrophys. 458, 21 (2006) The X-ray spectrum of the bursting atoll source 4U 1728-34 observed with INTEGRAL.

J/A+A/458/113/  w02-22.dat w08-33.dat w15-255.dat w3595-30.dat  (8.6 Mb)

Santiago B.X., Javiel S.C., Porto de Mello G.F. Astron. Astrophys. 458, 113 (2006) Multiband photometry towards the inner Galactic Bulge: extinction, atmospheric parameters and metallicities.

J/A+A/458/173/ gal/* pagb/* pec/* pn/* table1.dat table2.dat table3.dat table4.dat table5.dat table6.dat trans/* young/* (9.3 Mb) Suarez O., Garcia-Lario P., Manchado A., Manteiga M., Ulla A., Pottasch S.R. Astron. Astrophys., 458, 173-180 (2006) A spectroscopic atlas of post-AGB stars and planetary nebulae selected from the IRAS point source catalogue.

J/A+A/458/225/  GR290 field.jpg table2.dat  (789Kb)

Viotti R.F., Rossi C., Polcaro V.F., Montagni F., Gualandi R., Norci L. Astron. Astrophys. 458, 225 (2006) The present status of four luminous variables in M 33.

J/A+A/458/245/  appendix.ps appendix.ps chart.pdf table1.dat table7.dat  (941Kb)

La Palombara N., Mignani R.P., Hatziminaoglou E., Schirmer M., Bignami G.F., Caraveo P. Astron. Astrophys. 458, 245 (2006) XMM-Newton and ESO observations of the two Unidentified Gamma-ray Sources 3EG J0616-3310 and 3EG J1249-8330.

J/A+A/458/293/  table1.dat  (32Kb)

Paunzen E., Schnell A., Maitzen H.M. Astron. Astrophys. 458, 293 (2006) An empirical temperature calibration for the Δa photometric system. II. The A-type and mid F-type stars.

J/A+A/458/339/  fig/* table1.dat  (1.2 Mb)

Henden A. Munari U. Astron. Astrophys. 458, 339 (2006) UBV(RI)C photometric sequences for symbiotic stars. III.

J/A+A/458/341/  table3.dat  (42Kb)

Trachternach C., Bomans D.J., Haberzettl L., Dettmar R.-J. Astron. Astrophys. 458, 341 (2006) An optical search for Low Surface Brightness Galaxies in the Arecibo HI Strip Survey.

J/A+A/458/427/  linelist.dat table1.dat  (97Kb)

Janknecht E., Reimers D., Lopez S., Tytler D. Astron. Astrophys. 458, 427 (2006) The evolution of Lyman alpha absorbers in the redshift range 0.5 z 1.9.

J/A+A/458/461/  table1.dat table2.dat table6.dat  (28Kb)

Koehler R., Petr-Gotzens M.G., McCaughrean M.J., Bouvier J., Duchene G., Quirrenbach A., Zinnecker H.

Astron. Astrophys., 458, 461-476 (2006) Binary stars in the Orion Nebula Cluster.

J/A+A/458/533/  stars.dat tablea.dat  (34Kb)

Jimenez-Esteban F.M., Garcia-Lario P., Engels D., Manchado A. Astron. Astrophys. 458, 533 (2006) Near-IR variability properties of a selected sample of AGB stars.

J/A+A/458/541/  tablea1.dat tablea2.dat tablea3.dat tablea4.dat tablea5.dat tablea6.dat  (126Kb)

Beuermann K., Burwitz V., Rauch T. Astron. Astrophys. 458, 541 (2006) Establishing HZ43 A, Sirius B, and RXJ1856-37 as soft X-ray standards: a cross-calibration between the Chandra LETG+HRC-S, the EUVE spectrometer, and the ROSAT PSPC.

J/A+A/458/569/  table1.dat  (26Kb)

Wade G.A., Smith M.A., Bohlender D.A., Ryabchikova T.A., Bolton C.T., Lueftinger T., Landstreet J.D., Petit P., Strasser S., Blake M., Hill G.M. Astron. Astrophys., 458, 569-580 (2006) The magnetic Bp star 36 Lyncis. I. Magnetic and photospheric properties.

J/A+A/458/767/  tablea1.dat tablea2.dat  (26Kb)

Smriglio F., Dasgupta A.K., Boyle R.P. Astron. Astrophys., 458, 767-775 (2006) Photometric determination of the [Fe/H] ratio for the globular cluster M92.

J/A+A/458/789/  table1.dat table2.dat table3.dat  (78Kb)

Djupvik A.A., Andre P., Bontemps S., Motte F., Olofsson G., Gaalfalk M., Floren H.-G. Astron. Astrophys. 458, 789 (2006) A multi-wavelength census of star formation activity in the young embedded cluster around Serpens/G3-G6.

J/A+A/458/873/  table4.dat table23.dat  (15Kb)

Sousa S.G., Santos N.C., Israelian G., Mayor M., Monteiro M.J.P.F.G. Astron. Astrophys. 458, 873 (2006) Spectroscopic parameters for a sample of metal-rich solar-type stars.

J/A+A/458/895/  table2.dat  (8Kb)

de Medeiros J.R., Silva R.P., do Nascimento J.D.Jr, Canto Martins B.L., da Silva L., Melo C., Burnet M. Astron. Astrophys. 458, 895 (2006) A catalog of rotational and radial velocities for evolved stars. IV. Metal-poor stars.

J/A+A/458/997/  table1.dat table2.dat table7.dat tables.dat  (20Kb)

Santos N.C., Ecuvillon A., Israelinan G., Mayor M., Melo C., Queloz D., Udry S., Ribeiro J.P., Jorge S. Astron. Astrophys. 458, 997 (2006) Chemical abundances for the transiting planet host stars OGLE-TR-10, 56, 111, 113, 132 and TrES-1. Abundances in different galactic populations.

J/A+A/459/L1/  scldr4a.dat scldr4b.dat  (132Kb)

Einasto J., Einasto M., Saar E., Tago E., Liivamaegi L.J., Joeveer M., Suhhonenko I., Huetsi G., Jaaniste J., Heinaemaeki P., Mueller V., Knebe A., Tucker D. (5) Astron. Astrophys. 459, L1 (2006) Luminous superclusters: remnants from inflation?

J/A+A/459/L13/  galaxies.dat hii/*  (305Kb)

Bradley T.R., Knapen J.H., Beckman J.E., Folkes S.L. Astron. Astrophys. 459, L13 (2006) A composite HII region luminosity function in Hαof unprecedented statistical weight.

J/A+A/459/55/  table1.dat  (15Kb)

Bennert N., Jungwiert B., Komossa S., Haas M., Chini R. Astron. Astrophys. 459, 55 (2006) Size and properties of the narrow-line region in Seyfert-1 galaxies from spatially-resolved optical spectroscopy.

J/A+A/459/85/  refs.dat table1.dat  (20Kb)

Nagao T., Maiolino R., Marconi A. Astron. Astrophys., 459, 85-101 (2006) Gas metallicity diagnostics in star-forming galaxies.

J/A+A/459/137/ table2.dat table4.dat table5.dat table6.dat table7.dat table8.dat table9.dat table10.dat table11.dat tables.ps tables.tex (195Kb) Rivnius T., Baade D., Stefl S. Astron. Astrophys. 459, 137 (2006)

Bright Be shell stars.

J/A+A/459/199/  table2.dat tableb1.dat tablec1.dat tabled1.dat  (613Kb)

Argiroffi C., Favata F., Flaccomio E., Maggio A., Micela G., Peres G., Sciortino S. Astron. Astrophys., 459, 199 (2006) XMM-Newton survey of two Upper Scorpius regions.

J/A+A/459/307/  table1.dat table2.dat table3.dat  (29Kb)

Del Zanna G. Astron. Astrophys., 459, 307-316 (2006) Benchmarking atomic data for astrophysics: Fe XVIII.

J/A+A/459/317/  table2.dat  (26Kb)

Bruenken S., McCarthy M.C., Thaddeus P., Godfrey P.D., Brown R.D. Astron. Astrophys. 459, 317 (2006) Improved line frequencies for the nucleic acid base uracil for a radioastronomical search.

J/A+A/459/321/  table3.dat table4.dat table5a/* table5b/*  (57.3 Mb)

Vilardell F., Ribas I., Jordi C. Astron. Astrophys. 459, 321 (2006) Eclipsing binaries suitable for distance determination in the Andromeda galaxy.

J/A+A/459/391/  tablea1.dat tablea2.dat  (28Kb)

Schuberth Y., Richtler T., Dirsch B., Hilker M., Larsen S.S., Kissler-Patig M., Mebold U. Astron. Astrophys., 459, 391-406 (2006) Dynamics of the NGC 4636 globular cluster system. An extremely dark matter dominated galaxy?

J/A+A/459/423/  table4.dat  (65Kb)

Battaglia G., Tolstoy E., Helmi A., Irwin M.J., Letarte B., Jablonka P., Hill V., Venn K.A., Shetrone M.D., Arimoto N., Primas F., Kaufer A., Francois P., Szeifert T., Abel T., Sadakane K. Astron. Astrophys. 459, 423 (2006) The DART imaging and CaT survey of the Fornax dwarf spheroidal galaxy.

J/A+A/459/489/  notes.dat refs.dat table2.dat table3.dat  (171Kb)

De Marchi F., De Angeli F., Piotto G., Carraro G., Davies M.B. Astron. Astrophys. 459, 489 (2006) Search and analysis of blue straggler stars in open clusters.

J/A+A/459/557/  lcwhole.dat  (1.7 Mb)

Silvotti R., Bonanno A., Bernabei S., Fontaine G., Charpinet S., Leccia S., Kjeldsen H., Janulis R., Frasca A., Ostensen R., Kim S.-L., Park B.-G., Jiang X., Reed M.D., Patterson R.S., Gietzen K.M., Clark P.J., Wolf G.W., Lipkin Y., Formiggini L., Leibowitz E., Oswalt T.D., Rudkin M., Johnston K., Brassard P., Chayer P., Green E.M., Bergeron P. Astron. Astrophys. 459, 557 (2006) The rapidly pulsating subdwarf B star PG1325+101. I. Oscillation modes from multisite observations.

J/A+A/459/669/  table7.dat  (19Kb)

Kervella P., Thevenin F., Coude Du Foresto V., Mignard F. Astron. Astrophys., 459, 669-678 (2006) Deep imaging survey of the environment of αCentauri. I. Adaptive optics imaging of αCen B with VLT-NACO.

J/A+A/459/679/  table1.dat  (54Kb)

Adami C., Scheidegger R., Ulmer M., Durret F., Mazure A., West M.J., Conselice C.J., Gregg M., Kasun S., Pello R., Picat J.P. Astron. Astrophys., 459, 679-692 (2006) A deep wide survey of faint low surface brightness galaxies in the direction of the Coma cluster of galaxies.

J/A+A/459/693/  table2.dat  (18Kb)

Akylas A., Georgantopoulos I., Georgakakis A., Kitsionas S., Hatziminaoglou E. Astron. Astrophys., 459, 693-701 (2006) XMM-Newton and Chandra measurements of the AGN intrinsic absorption: dependence on luminosity and redshift.

J/A+A/459/777/  tablea1.dat tablea2.dat  (20Kb)

Ramsay G., Wu K. Astron. Astrophys., 459, 777-781 (2006) A search for stellar X-ray sources in the Sagittarius and Carina dwarf galaxies. I. X-ray observations.

J/A+A/459/849/  table2.dat  (17Kb)

Koubsky P., Harmanec P., Yang S., Netolicky M., Skoda P., Slechta M., Korcakova D. Astron. Astrophys. 459, 849 (2006)

Properties and nature of Be stars. 25. A new orbital solution and the nature of a peculiar emission-line binary upsilon Sgr.

J/A+A/459/871/  table1.dat  (60Kb)

Paunzen E., Maitzen H.M., Pintado O.I., Claret A., Iliev I.K., Netopil M. Astron. Astrophys. 459, 871 (2006) Chemically peculiar stars in the Large Magellanic Cloud

J/A+A/459/875/  obs.dat sp/*  (31.6 Mb)

Ederoclite A., Mason E., Della Valle M., Gilmozzi R., Williams R.E., Germany L., Saviane I., Matteucci F., Schaefer B.E., Walter F., Rudy R.J., Lynch D., Mazuk S., Venturini C.C., Puetter R.C., Perry R.B., Liller W., Rotter A. Astron. Astrophys. 459, 875 (2006) Early spectral evolution of Nova Sgr 2004 (V5114 Sgr).

J/A+A/459/981/  sxy2p03.dat xp03.dat yp03.dat  (231Kb)

Wallace P.T., Capitaine N. Astron. Astrophys. 459, 981 (2006) Precession-nutation procedures consistent with IAU 2006 resolutions.

J/A+A/460/133/  table2.dat table3.dat  (45Kb)

Damiani F., Micela G., Sciortino S., Huelamo N., Moitinho A., Harnden F.R.Jr., Murray S.S. Astron. Astrophys. 460, 133 (2006) The young star cluster NGC 2362: low-mass population and initial mass function from a Chandra X-ray observation.

J/A+A/460/331/  table1.dat table2.dat table3.dat table4.dat tables.ps tables.tex  (27.2 Mb)

Aggarwal K.M., Keenan F.P., Kato T., Murakami I. Astron. Astrophys. 460, 331 (2006) Energy levels and radiative rates for transitions in Fe IX.

J/A+A/460/695/  refs.dat table3.dat table4.dat  (417Kb)

Torres C.A.O., Quast G.R., da Silva L., de la Reza R., Melo C.H.F., Sterzik M. Astron. Astrophys. 460, 695 (2006) Search for Associations Containing Young stars (SACY). I. Sample + Searching Method.

J/A+A/460/783/  table2.dat  (7Kb)

Beuermann K. Astron. Astrophys. 460, 783 (2006) Barnes-Evans relations for dwarfs with an application to the determination of distances to cataclysmic variables.

J/A+A/460/959/  table1.dat table2.dat table3.dat table4.dat tables.ps tables.tex  (2.1 Mb)

Aggarwal K.M., Keenan F.P. Astron. Astrophys. 460, 959 (2006) Energy levels and radiative rates for transitions in Ni XIX.

J/A+A/461/11/ fig3/* fig7/* fig8/* table3.dat table5.dat table6.dat table7.dat table8.dat (52.1 Mb) Urquhart J.S., Busfield A.L., Hoare M.G., Lumsden S.L., Clarke A.J., Moore T.J.T., Mottram J.C., Oudmaijer R.D. Astron. Astrophys., 461, 11-23 (2007) The RMS Survey: Radio observations of candidate massive YSOs in the southern hemisphere.

J/A+A/461/81/  table3.dat  (18Kb)

Olsen L.F., Benoist C., Cappi A., Maurogordato S., Mazure A., Slezak E., Adami C., Ferrari C., Martel F. Astron. Astrophys. 461, 81 (2007) Galaxy clusters in the CFHTLS. First matched filter candidate catalogue of the Deep fields.

J/A+A/461/183/  phot/* stars.dat  (348Kb)

Grankin K.N., Melnikov S.Y., Bouvier J., Herbst W., Shevchenko V.S. Astron. Astrophys. 461, 183 (2007) Results of the ROTOR-program. I. The long-term photometric variability of classical T Tauri stars.

J/A+A/461/197/  fig/* sources.dat  (4.2 Mb)

Wang Y., Wu Y., Zhang Q., Mao R.-Q., Miller M. Astron. Astrophys. 461, 197 (2007) VLA NH3 observations of regions of massive star formation in protostellar cores.

J/A+A/461/397/  tablea1.dat  (17Kb)

Popesso P., Biviano A., Boehringer H., Romaniello M. Astron. Astrophys., 461, 397-410 (2007)

RASS-SDSS galaxy cluster survey. V. The X-ray-underluminous Abell clusters.

J/A+A/461/509/  ic2391.dat table1.dat table3.dat table5.dat table6.dat  (420Kb)

Platais I., Melo C., Mermilliod J.-C., Kozhurina-Platais V., Fulbright J.P., Mendez R.A., Altmann M., Sperauskas J. Astron. Astrophys. 461, 509 (2007) WIYN Open Cluster Study. XXVI. Improved kinematic membership and spectroscopy of IC 2391.

J/A+A/461/613/  table1.dat table2.dat  (32Kb)

Szabados L., Kiss L.L., Derekas A. Astron. Astrophys. 461, 613 (2007) The anomalous Cepheid XZ Ceti

J/A+A/461/631/ (3Kb) Negueruela I., Schurch M.P.E. Astron. Astrophys., 461, 631-639 (2007) A search for counterparts to massive X-ray binaries using photometric catalogues.

J/A+A/461/669/  table4.dat  (16Kb)

Robrade J., Schmitt J.H.M.M. Astron. Astrophys., 461, 669-678 (2007) A deep XMM-Newton X-ray observation of the Chamaeleon I dark cloud.

J/A+A/461/789/ plots/* table3.dat table4.dat table5.dat table6.dat table7.dat table8.dat table9.dat (8.9 Mb) Cremonese G., Capria M.T., De Sanctis M.C. Astron. Astrophys. 461, 789 (2007) Catalog of the emission lines in the visible spectrum of comet 153P/Ikeya-Zhang.

J/A+A/461/823/  table1.dat tablea1.dat tablea2.dat tablea3.dat tablea4.dat tablea5.dat  (24Kb)

Venemans B.P., Roettgering H.J.A., Miley G.K., Van Breugel W.J.M., De Breuck C., Kurk J.D., Pentericci L., Stanford S.A., Overzier R.A., Croft S., Ford H. Astron. Astrophys., 461, 823-845 (2007) Protoclusters associated with z 2 radio galaxies. I. Characteristics of high redshift protoclusters.

J/A+A/461/949/  table2.dat  (9Kb)

Ortiz R., Roman-Lopes A., Abraham Z. Astron. Astrophys. 461, 949 (2007) The young stellar cluster associated with IRAS 09149-4743.

J/A+A/461/977/  table1.dat table2.dat  (19Kb)

Micela G., Affer L., Favata F., Henry J.P., Gioia I., Mullis C.R., Sanz-Forcada J., Sciortino D. Astron. Astrophys. 461, 977 (2007) The stellar population of the Rosat North Ecliptic survey.

J/A+A/461/1065/  table11.dat table12.dat table13.dat table14.dat  (200Kb)

Clausen J.V., Helt B.E., Gimenez A., Vaz L.P.R., Garcia J.M., Olsen E.H., Southworth J. Astron. Astrophys. 461, 1065 (2007) Four-colour photometry of eclipsing binaries. XL. uvby light curves for the B-type systems DW Carinae, BF Centauri, AC Velorum, and NSV 5783.

J/A+A/461/1077/  table1.dat  (8Kb)

Southworth J., Clausen J.V. Astron. Astrophys. 461, 1077 (2007) Absolute dimensions of eclipsing binaries. XXIV. The Be star system DW Carinae, a member of the open cluster Collinder 228.

J/A+A/462/L17/  table1.dat  (10Kb)

Alves J., Lombardi M., Lada C. Astron. Astrophys. 462, L17 (2007) The mass function of dense molecular cores and the origin of the IMF.

J/A+A/462/L9/  rv.dat  (20Kb)

Scarpa R., Marconi G., Gilmozzi R., Carraro G. Astron. Astrophys. 462, L9 (2007) Using globular clusters to test gravity in the weak acceleration regime: NGC 7099.

J/A+A/462/81/  table4.dat table5.dat  (25Kb)

Chanial P., Flores H., Guiderdoni B., Elbaz D., Hammer F., Vigroux L. Astron. Astrophys., 462, 81-91 (2007) The infrared compactness-temperature relation for quiescent and starburst galaxies.

J/A+A/462/107/  table2.dat table3.dat table6.dat  (74Kb)

Fagiolini M., Raimondo G., Degl’Innocenti S. Astron. Astrophys. 462, 107 (2007) Monte Carlo simulations of metal-poor star clusters.

J/A+A/462/123/  table1.dat table3.dat  (81Kb)

Prisinzano L., Damiani F., Micela G., Pillitteri I. Astron. Astrophys. 462, 123 (2007) VLT/Flames observations of the star forming region NGC 6530.

J/A+A/462/157/  table1.dat  (46Kb)

Paunzen E., Netopil M., Zwintz K. Astron. Astrophys., 462, 157-162 (2007) Investigating star formation in the young open cluster NGC 6383.

J/A+A/462/245/  catalog.dat  (5.7 Mb)

Guarcello M.G., Prisinzano L., Micela G., Damiani F., Peres G., Sciortino S. Astron. Astrophys., 462, 245-255 (2007) Correlation between the spatial distribution of circumstellar disks and massive stars in the open cluster NGC 6611. Compiled catalog and cluster parameters.

J/A+A/462/449/  table1.dat table2.dat  (20Kb)

Branchesi M., Gioia I.M., Fanti C., Fanti R., Cappelluti N. Astron. Astrophys., 462, 449-458 (2007) Chandra point-source counts in distant galaxy clusters.

J/A+A/462/507/  table2.dat table3.dat  (117Kb)

Lisenfeld U., Verdes-Montenegro L., Sulentic J., Leon S., Espada D., Bergond G., Garcia E., Sabater J., Santander-Vela J.D., Verley S. Astron. Astrophys. 462, 507 (2007) The AMIGA sample of isolated galaxies. III. IRAS data and infrared diagnostics.

J/A+A/462/535/  table1.dat  (64Kb)

Yin S.Y., Liang Y.C., Hammer F., Brinchmann J., Zhang B., Deng L.C., Flores H. Astron. Astrophys., 462, 535-546 (2007) Empirical strong-line oxygen abundance calibrations from galaxies with electron-temperature measurements.

J/A+A/462/553/  cat/* deep1.dat deep2.dat deep3.dat desccat.dat fields.dat filter/*  (690.5 Mb)

Mignano A., Miralles J.-M., Da Costa L., Olsen L.F., Prandoni I., Arnouts S., Benoist C., Madejsky R., Slijkhuis R., Zaggia S. Astron. Astrophys., 462, 553-563 (2007) ESO imaging survey: optical deep public survey.

J/A+A/462/591/ (3Kb) Netopil M., Paunzen E., Maitzen H.M., Pintado O.I., Claret A., Miranda L.F., Iliev I.K., Casanova V. Astron. Astrophys., 462, 591-597 (2007) CCD photometric search for peculiar stars in open clusters. VIII. King 21, NGC 3293, NGC 5999, NGC 6802, NGC 6830, Ruprecht 44, Ruprecht 115, and Ruprecht 120.

J/A+A/462/599/  photom.dat table23.dat  (28Kb)

Testa V., Marconi M., Musella I., Ripepi V., Dallora M., Ferraro F.R., Mucciarelli A., Mateo M., Cote P. Astron. Astrophys. 462, 599 (2007) IR photometry of Cepheids in the LMC clusters NGC 1866 and NGC 2031.

J/A+A/462/621/  table1.dat  (9Kb)

Vergne M.M., Feinstein C., Martinez R. Astron. Astrophys. 462, 621 (2007) Optical multicolor polarization observations in the region of the open cluster NGC 5749.

J/A+A/462/683/  table1.dat table3.dat  (44Kb)

Martayan C., Fremat Y., Hubert A.-M., Floquet M., Zorec J., Neiner C. Astron. Astrophys., 462, 683-694 (2007) Effects of metallicity, star-formation conditions, and evolution in B and Be stars. II. Small Magellanic Cloud, field of NGC 330.

J/A+A/462/795/  table5.dat  (1.2 Mb)

Zwitter T., Mignard F., Crifo F. Astron. Astrophys. 462, 795 (2007) Asteroids as radial velocity and resolving power standards for medium and high resolution spectroscopy.

J/A+A/462/811/  scl2dfna.dat scl2dfnb.dat scl2dfsa.dat scl2dfsb.dat  (74Kb)

Einasto J., Einasto M., Tago E., Saar E., Hutsi G., Joeveer M., Liivamagi L.J., Suhhonenko I., Jaaniste J., Heinamaki P., Muller V., Knebe A., Tucker D. Astron. Astrophys. 462, 811 (2007) Superclusters of galaxies from the 2dF redshift survey. I. The catalogue.

J/A+A/462/1081/  table2.dat table3.dat  (41Kb)

Reig P., Larionov V., Negueruela I., Arkharov A.A., Kudryavtseva N.A. Astron. Astrophys. 462, 1081 (2007) The Be/X-ray transient 4U 0115+63/V635 Cassiopeiae. III. Quasi-cyclic variability.

J/A+A/463/97/  refs.dat table2.dat table4.dat  (27Kb)

Labiano A., Barthel P.D., O’dea C.P., De Vries W.H., Perez I., Baum S.A. Astron. Astrophys., 463, 97-104 (2007) GPS radio sources: new optical observations and an updated master list.

J/A+A/463/175/  /2mass.dat irac.dat  (19Kb)

Vig S., Ghosh S.K., Ojha D.K., Verma R.P. Astron. Astrophys., 463, 175-186 (2007) Infrared study of the southern Galactic star-forming region associated with IRAS 14416-5937.

J/A+A/463/233/  table1.dat table2.dat table3.dat  (44Kb)

Ak H., Chadima P., Harmanec P., Demircan O., Yang S., Koubsky P., Skoda P., Slechta M., Wolf M., Bozic H., Ruzdjak D., Sudar D. Astron. Astrophys. 463, 233 (2007) New findings supporting the presence of a thick disc and bipolar jets in the beta Lyrae system.

J/A+A/463/243/  stars.dat table2a.dat table2b.dat table3a.dat table3b.dat  (53Kb)

De Cat P., Briquet M., Aerts C., Goossens K., Saesen S., Cuypers J., Yakut K., Scuflaire R., Dupret M.-A., Uytterhoeven K., Van Winckel H., Raskin G., Davignon G., Le Guillou L., Van Malderen R., Reyniers M., Acke B., De Meester W., Vanautgaerden J., Vandenbussche B., Verhoelst T., Waelkens C., Deroo P., Reyniers K., Ausseloos M., Broeders E., Daszynska-Daskiewicz J., Debosscher J., De Ruyter S., Lefever K., Decin G., Kolenberg K., Mazumdar A., Van Kerckhoven C., De Ridder J., Drummond R., Barban C., Vanhollebeke E., Maas T., Decin L. Astron. Astrophys. 463, 243 (2007) Long term photometric monitoring with the Mercator telescope: Frequencies and mode identification of variable O-B stars.

J/A+A/463/265/  objects.dat table.dat  (92Kb)

Girard P., Koeppen J., Acker A. Astron. Astrophys. 463, 265 (2007) Chemical compositions and plasma parameters of planetary nebulae with Wolf-Rayet and wels type central stars.

J/A+A/463/275/  table5.dat table6.dat  (20Kb)

Giardino G., Favata F., Micela G., Sciortino S., Winston E. Astron. Astrophys., 463, 275-288 (2007) The onset of X-ray emission in young stellar objects. A Chandra observation of the Serpens star-forming region.

J/A+A/463/399/  table1.dat table2.dat table3.dat tables.ps tables.tex  (4.5 Mb)

Aggarwal K.M., Keenan F.P. Astron. Astrophys. 463, 399 (2007) Energy levels and radiative rates for inner shell transitions of Fe XVI.

J/A+A/463/455/ galaxies.dat table5.dat table6.dat table7.dat table9.dat table10.dat table11.dat table12.dat (147Kb) Annibali F., Bressan A., Rampazzo R., Zeilinger W.W., Danese L. Astron. Astrophys. 463, 455 (2007) Nearby early-type galaxies with ionized gas. III. Analysis of line-strength indices with new stellar population models.

J/A+A/463/503/  table1.dat  (17Kb)

Mieske S., Hilker M., Infante L., Mendes de Oliveira C. Astron. Astrophys. 463, 503 (2007) The early-type dwarf galaxy population in the Fornax cluster.

J/A+A/463/519/  table.dat  (71Kb)

Bondi M., Ciliegi P., Venturi T., Dallacasa D., Bardelli S., Zucca E., Athreya R.M., Gregorini L., Zanichelli A., Le Fevre O., Contini T., Garilli B., Iovino A., Temporin S., Vergani D. Astron. Astrophys. 463, 519 (2007)

The VVDS-VLA Deep Field: III. GMRT observations at 610 MHz and the radio spectral index properties of the sub-mJy population.

J/A+A/463/579/  table3.dat  (17Kb)

Groenewegen M.A.T., Decin L., Salaris M., De Cat P. Astron. Astrophys. 463, 579 (2007) The Pleiades eclipsing binary HD 23642 revisited.

J/A+A/463/589/  isis/* psf/* rv/* rvp/* stars.dat varstars.dat  (2.7 Mb)

Szekely P., Kiss L.L., Jackson R., Derekas A., Csak B., Szatmary K. Astron. Astrophys. 463, 589 (2007) RR Lyrae stars in the southern globular cluster NGC 362.

J/A+A/463/671/  table4.dat  (60Kb)

Royer F., Zorec J., Gomez A.E. Astron. Astrophys. 463, 671 (2007) Rotational velocities of A-type stars. III. Velocity distributions.

J/A+A/463/703/  table1.dat  (8Kb)

Gromadzki M., Mikolajewska J., Whitelock P.A., Marang F. Astron. Astrophys. 463, 703 (2007) On the nature of the cool component of MWC 560.

J/A+A/463/783/  eoc3.dat  (1.3 Mb)

Vondrak J., Stefka V. Astron. Astrophys. 463, 783 (2007) Combined astrometric catalogue EOC-3.

J/A+A/463/789/  notes.dat refs.dat table8.dat table9.dat  (244Kb)

Ahumada J.A., Lapasset E. Astron. Astrophys. 463, 789 (2007) New catalogue of blue stragglers in open clusters.

J/A+A/463/981/  table4.dat table23.dat  (120Kb)

Rauw G., Manfroid J., Gosset E., Naze Y., Sana H., De Becker M., Foellmi C., Moffat A.F.J. Astron. Astrophys. 463, 981 (2007) Early-type stars in the core of the young open cluster Westerlund 2.

J/A+A/463/1017/  aatau.dat  (9Kb)

Bouvier J., Alencar S.H.P., Boutelier T., Dougados C., Balog Z., Grankin K., Hodgkin S.T., Ibrahimov M.A., Kun M., Magakian T.Yu., Pinte C. Astron. Astrophys. 463, 1017 (2007) Magnetospheric accretion-ejection processes in the classical T Tauri star AA Tau.

J/A+A/463/1029/  dc303.dat tpn.dat  (171Kb)

Kainulainen J., Lehtinen K., Vaeisaenen P., Bronfman L., Knude J. Astron. Astrophys. 463, 1029 (2007) A comparison of density structures of a star forming and a non-star forming globule. DC303-14.8 and Thumbprint Nebula.

J/A+A/463/1053/  table4.dat  (55Kb)

Sterken C., Vogt N., Uemura M., Tuvikene T. Astron. Astrophys. 463, 1053 (2007) WX Ceti: a closer look at its behaviour in quiescence and outburst.

J/A+A/463/1061/  table2a.dat table2b.dat table2c.dat table2d.dat  (98Kb)

Harmanec P., Mayer P., Prsa A., Bozic H., Eenens P., Guinan E.F., McCook G., Koubsky P., Ruzdjak D., Engle S., Sudar D., Skoda P., Slechta M., Wolf M., Yang S. Astron. Astrophys. 463, 1061 (2007) V379 Cep: a quadruple system of two binaries.

J/A+A/463/1227/  appenb.dat appenc.dat append.dat refs.dat  (378Kb)

Konyves V., Kiss Cs., Moor A., Kiss Z.T., Toth L.V. Astron. Astrophys. 463, 1227 (2007) Catalogue of far-infrared loops in the Galaxy.

J/A+A/464/L21/  table1.dat table2.dat table3.dat  (19Kb)

Bergond G., Athanassoula E., Leon S., Balkowski C., Cayatte V., Chemin L., Guzman R., Meylan G., Prugniel P. Astron. Astrophys., 464, L21-L24 (2007) Globular clusters and dwarf galaxies in Fornax. I. Kinematics in the cluster core from multi-object spec-

troscopy.

J/A+A/464/201/  appendix.dat table1.dat  (49Kb)

Monaco L., Bellazzini M., Bonifacio P., Buzzoni A., Ferraro F.R., Marconi G., Sbordone L., Zaggia S. Astron. Astrophys., 464, 201-209 (2007) High-resolution spectroscopy of RGB stars in the Sagittarius streams. I. Radial velocities and chemical abundances.

J/A+A/464/211/  table1.dat table2.dat table3.dat  (254Kb)

Albacete Colombo J.F., Flaccomio E., Micela G., Sciortino S., Damiani F. Astron. Astrophys., 464, 211-227 (2007) Unveiling the Cygnus OB2 stellar population with Chandra.

J/A+A/464/263/  table3.dat  (45Kb)

Bozic H., Wolf M., Harmanec P., Prsa A., Percy J.R., Ruzdjak D., Sudar D., Slechta M., Ak H., Eenens P. Astron. Astrophys. 464, 263 (2007) HD 143418: An unusual light variable and a double-lined spectroscopic binary with a CP primary.

J/A+A/464/373/  table1.dat  (490Kb)

Kervella P., Thevenin F. Astron. Astrophys. 464, 373 (2007) Deep imaging survey of the environment of Alpha Centauri II. CCD imaging with the NTT-SUSI2 camera.

J/A+A/464/377/  table2.dat  (162Kb)

Frankowski A., Jancart S., Jorissen A. Astron. Astrophys. 464, 377 (2007) Proper-motion binaries in the Hipparcos Catalogue Comparison with radial velocity data.

J/A+A/464/495/  tables.dat  (15Kb)

Mora M.D., Larsen S.S., Kissler-Patig M. Astron. Astrophys., 464, 495-505 (2007) Imaging of star clusters in unperturbed spiral galaxies with the Advanced Camera for Surveys. I. The low luminosity galaxy NGC 45.

J/A+A/464/573/  table1.dat table2.dat table3.dat table4.dat table5.dat  (1.5 Mb)

Carraro G., Costa E. Astron. Astrophys. 464, 573 (2007) Photometry of the five marginally studied open clusters Collinder 74, Berkeley 27, Haffner 8, NGC 2509, and VdB-Hagen 4.

J/A+A/464/581/  table2.dat table4.dat tablea1.dat tablea2.dat tablea3.dat  (31Kb)

Kouwenhoven M.B.N., Brown A.G.A., Kaper L. Astron. Astrophys. 464, 581 (2007) A brown dwarf desert for intermediate mass stars in Scorpius OB2?

J/A+A/464/641/  table4.dat table5.dat table6.dat  (34Kb)

Lampens P., Strigachev A., Duval D. Astron. Astrophys. 464, 641 (2007) Multicolour CCD measurements of visual double and multiple stars. III.

J/A+A/464/885/  table3a.dat table3b.dat table4.dat  (147Kb)

Guseva N.G., Izotov Y.I., Papaderos P., Fricke K.J. Astron. Astrophys. 464, 885 (2007) Balmer jump temperature determination in a large sample of low-metallicity H II regions.

J/A+A/464/909/  table4.dat  (9Kb)

Brand J., Wouterloot J.G.A. Astron. Astrophys., 464, 909-920 (2007) A star cluster at the edge of the Galaxy.

J/A+A/464/927/  table2.dat table3.dat table5.dat tablea.dat  (342Kb)

Carreta E., Bragaglia A., Gratton R.G., Lucatello S., Momany Y. Astron. Astrophys. 464, 927 (2007) Na-O Anticorrelation and HB. II. The Na-O anticorrelation in the globular cluster NGC 6752

J/A+A/464/939/  table2.dat table3.dat table5.dat tablea.dat  (206Kb)

Carretta E., Bragaglia A., Gratton R.G., Catanzaro G., Leone F., Sabbi E., Cassisi S., Claudi R., D’Antona F., Francois P., James G., Piotto G. Astron. Astrophys. 464, 939 (2007) Na-O Anticorrelation and HB. IV. Detection of He-rich and He-poor stellar populations in the globular cluster NGC 6218.

J/A+A/464/953/  table2.dat table3.dat  (99Kb)

Gratton R.G., LucatelloS., Bragaglia A., Carretta E., Cassisi S., Momany Y., Pancino E., Valenti E., Caloi V., Claudi R., D’Antona F., Desidera S., Francois P., James G., Moehler S., Ortolani S., Pasquini L., Piotto G., Recio-Blanco A. Astron. Astrophys. 464, 953 (2007) Na-O anticorrelation and horizontal branches. V. The Na-O anticorrelation in NGC 6441 from Giraffe spectra.

J/A+A/464/967/  table2.dat table3.dat  (103Kb)

Carreta E., Bragaglia A., Gratton R.G., Momany Y., Recio-Blanco A., Cassisi S., Francois P., James G., Lucatello S., Moehler S. Astron. Astrophys. 464, 967 (2007) Na-O Anticorrelation and HB. VI. The chemical composition of the peculiar bulge globular cluster NGC 6388.

J/A+A/465/71/  sp/* spectra.dat table2.dat table3.dat table5.dat  (8.6 Mb)

Theureau G., Hanski M., Coudreau N., Hallet N., Martin J.-M. Astron. Astrophys. 465, 71 (2007) Kinematics of the Local Universe. XIII. 21-cm line measurements of 452 galaxies with the Nancay radiotelescope, JHK Tully-Fisher relation and preliminary maps of the peculiar velocity field.

J/A+A/465/185/  table2.dat table3.dat  (26Kb)

Sestito P., Randich S., Bragaglia A. Astron. Astrophys. 465, 185 (2007) Element abundances in the metal rich open cluster NGC 6253.

J/A+A/465/249/  table1.dat  (10Kb)

Althaus L.G., Garcia-Berro E., Isern J., Corsico A., Rohrmann R. Astron. Astrophys. 465, 249 (2007) The age and colors of massive white dwarf stars.

J/A+A/465/257/  table4.dat  (13Kb)

Tokovinin A.A., Gorynya N.G. Astron. Astrophys. 465, 257 (2007) New spectroscopic components in multiple systems. V.

J/A+A/465/271/  table4.dat table5.dat table6.dat  (119Kb)

Ramirez, I., Allende Prieto, C., Lambert, D.L. Astron. Astrophys. 465, 271 (2007) Oxygen abundances in nearby stars. Clues to the formation and evolution of the Galactic disk.

J/A+A/465/331/  lc/* objects.dat  (451Kb)

Durech J., Kaasalainen M., Marciniak A., Allen W.H., Behrend R., Bembrick C., Bennett T., Bernasconi L., Berthier J., Bolt G., Boroumand S., Crespo da Silva L., Crippa R., Crow M., Durkee R., Dymock R., Fagas M., Fauerbach M., Fauvaud S., Frey M., Goncalves R., Hirsch R., Jardine D., Kaminski K., Koff R., Kwiatkowski T., Lopez A., Manzini F., Michalowski T., Pacheco R., Pan M., Pilcher F., Poncy R., Pray D., Pych W., Roy R., Santacana G., Slivan S., Sposetti S., Stephens R., Warner B., Wolf M. Astron. Astrophys. 465, 331 (2007) Physical models of ten asteroids from observers collaboration network.

J/A+A/465/357/  table2.dat table6.dat  (29Kb)

Faria D., Feltzing S., Lundstroem I., Gilmore G., Wahlgren G.M., Ardeberg A., Linde P. Astron. Astrophys., 465, 357-373 (2007) The usage of Stroemgren photometry in studies of local group dwarf spheroidal galaxies. Application to Draco: a new catalogue of Draco members and a study of the metallicity distribution function and radial gradients.

J/A+A/465/375/  refs.dat tablea1.dat tableb2.dat  (187Kb)

Pietsch W., Haberl F., Sala G., Stiele H., Hornoch K., Riffeser A., Fliri J., Bender R., Buhler S., Burwitz V., Greiner J., Seitz S. Astron. Astrophys. 465, 375 (2007) X-ray monitoring of optical novae in M 31 from July 2004 to February 2005.

J/A+A/465/469/  fig/* table1.dat  (355Kb)

Sacuto S., Chesneau O., Vannier M., Cruzalebes P. Astron. Astrophys. 465, 469 (2007) A compact dust shell in the symbiotic system HM Sge.

J/A+A/465/481/  table2.dat  (8Kb)

Santander-Garcia M., Corradi R.L.M., Whitelock P.A., Munari U., Mampaso A., Marang F., Boffi F., Livio M. Astron. Astrophys. 465, 481 (2007) HST and VLT observations of the Symbiotic Star Hen 2-147: Its Nebular Dynamics, its Mira Variable and its

Distance.

J/A+A/465/799/  abund.dat ew.dat table1.dat  (171Kb)

Lecureur A., Hill V., Zoccali M., Barbuy B., Gomez A., Minniti D., Ortolani S., Renzini A. Astron. Astrophys., 465, 799-814 (2007) Oxygen, sodium, magnesium, and aluminium as tracers of the Galactic Bulge formation.

J/A+A/465/815/  refs.dat table1.dat tablea.dat  (97Kb)

Sbordone L., Bonifacio P., Buonanno R., Marconi G., Monaco L., Zaggia S. Astron. Astrophys., 465, 815-824 (2007) The exotic chemical composition of the Sagittarius dwarf spheroidal galaxy.

J/A+A/465/943/  photjd/* photph/* rv/* stars.dat  (56Kb)

Szalai T., Kiss L.L., Meszaros Sz., Vinko J., Csizmadia Sz. Astron. Astrophys. 465, 943 (2007) Physical parameters and multiplicity of five southern close eclipsing binaries.

J/A+A/465/965/  stars.dat table1.dat table2.dat table3.dat table4.dat  (2.3 Mb)

Arentoft T., De Ridder J., Grundahl F., Glowienka L., Waelkens C., Dupret M.-A., Grigahcene A., Lefever K., Jensen H. R., Reyniers M., Frandsen S., Kjeldsen H. Astron. Astrophys. 465, 965 (2007) Oscillating blue stragglers, gamma Doradus stars and eclipsing binaries in the open cluster NGC 2506.

J/A+A/465/993/  refs.dat sp/* table1.dat  (337Kb)

Groh J.H., Damineli A., Jablonski F. Astron. Astrophys. 465, 993 (2007) Spectral atlas of massive stars around He I 10830Å.

J/A+A/465/1099/  table2.dat  (61Kb)

Ravikumar C.D., Puech M., Flores H., Proust D., Hammer F., Lehnert M., Rawat A., Amram P., Balkowski C., Burgarella D., Cassata P., Cesarsky C., Cimatti A., Combes F., Daddi E., Dannerbauer H., di Serego Alighieri S., Elbaz D., Guiderdoni B., Kembhavi A., Liang Y.C., Pozzetti L., Vergani D., Vernet J., Wozniak H., Zheng X.Z. Astron. Astrophys. 465, 1099 (2007) New spectroscopic redshifts from the CDFS and a test of the cosmological relevance of the GOODS-South field.

J/A+A/466/L1/  fig/* table1.dat  (48Kb)

Leao I., de Laverny P., Chesneau O., Mekarnia D., De Medeiros J.R. Astron. Astrophys. 466, L1 (2007) A snapshot of the inner dusty regions of a R CrB-type variable.

J/A+A/466/31/  table1.dat table3.dat table4.dat  (33Kb)

Cocchia F., Fiore F., Vignali C., Mignoli M., Brusa M., Comastri A., Feruglio C., Bald A., Carangelo N., Ciliegi P., D’Elia V., La Franca F., Maiolino R., Matt G., Molendi S., Perola G.C., Puccetti S. Astron. Astrophys. 466, 31 (2007) The HELLAS2XMM survey: VIII. Optical identifications of the extended sample.

J/A+A/466/41/  notes.dat opt.dat xray.dat  (81Kb)

Krumpe M., Lamer G., Schwope A.D., Wagner S., Zamorani G., Mignoli M., Staubert R., Wisotzki L., Hasinger G Astron. Astrophys. 466, 41 (2007) The XMM-Newton survey in the Marano Field. I. The X-ray data and optical follow-up.

J/A+A/466/63/  refs.dat table4.dat  (22Kb)

Wu Z., Jiang D.R., Gu M., Liu Y. Astron. Astrophys., 466, 63-73 (2007) VLBI observations of seven BL Lacertae objects from RGB sample.

J/A+A/466/75/  images.dat images/* table1.dat table2.dat  (99.5 Mb)

Pena M., Richer M.G., Stasinska G. Astron. Astrophys. 466, 75 (2007) A catalog of planetary nebulae and HII regions in NGC 3109.

J/A+A/466/137/  w2phot.dat  (374Kb)

Ascenso J., Alves J., Beletsky Y., Lago M.T.V.T. Astron. Astrophys. 466, 137 (2007) Near-infrared imaging of Galactic massive clusters: Westerlund 2.

J/A+A/466/157/  sp/* table2.dat  (123Kb)

Soto M., Minniti D., Rejkuba M.

Astron. Astrophys. 466, 157 (2007) Spectroscopic and light curve characterization of Bulge Microlensing Events.

J/A+A/466/181/  table1.dat  (69Kb)

Carraro G., Zinn R., Moni Bidin C. Astron. Astrophys. 466, 181 (2007) Whiting 1: the youngest globular cluster associated with the Sagittarius dwarf spheroidal galaxy.

J/A+A/466/255/  table6.dat table7.dat table8.dat  (1.1 Mb)

Demyk K., Maeder H., Tercero B., Cernicharo J., Demaison J., Magules L., Wegner M., Keipert S., Sheng M. Astron. Astrophys. 466, 255 (2007) Isotopic ethyl cyanide ^{13}CH3CH2CN, CH3^{13}CH2CN, and CH3CH2^{13}CN: laboratory rotational spectrum and detection in Orion.

J/A+A/466/277/  stars.dat table3.dat table4.dat table5.dat table6.dat  (119Kb)

Hunter I., Dufton P.L., Smartt S.J., Ryans R.S.I., Evans C.J., Lennon D.J., Trundle C., Hubeny I., Lanz T. Astron. Astrophys., 466, 277-300 (2007) The VLT-FLAMES survey of massive stars: surface chemical compositions of B-type stars in the Magellanic Clouds.

J/A+A/466/481/  table1.dat table2.dat table3.dat table4.dat  (51Kb)

Schroeder A.C., Mamon G.A., Kraan-Korteweg R.C., Woudt P.A. Astron. Astrophys., 466, 481-499 (2007) The highly obscured region around PKS 1343-601. I. Galactic interstellar extinctions using DENIS galaxy colours.

J/A+A/466/501/  table1.dat  (36Kb)

Valcheva A.T., Ivanov V.D., Ovcharov E.P., Nedialkov P.L. Astron. Astrophys., 466, 501-507 (2007) Carbon stars and C/M ratio in the WLM dwarf irregular galaxy.

J/A+A/466/627/  table2.dat  (178Kb)

Hu Q., Wu C., Wu X.-B. Astron. Astrophys. 466, 627 (2007) The mass and luminosity functions and the formation rate of DA white dwarfs in the Sloan Digital Sky Survey.

J/A+A/466/641/  table13.dat table13.tex  (131Kb)

Izzard R.G., Lugaro M., Karakas A.I., Iliadis C., van Raai M. Astron. Astrophys. 466, 641 (2007) Reaction rate uncertainties and the operation of the NeNa and MgAl chains during HBB in intermediate-mass AGB stars.

J/A+A/466/743/  o132 ntt.dat o132 vlt.dat  (20Kb)

Gillon M., Pont F., Moutou C., Santos N. C., Bouchy F., Hartman J. D., Mayor M., Melo C., Queloz D., Udry S., Magain P. Astron. Astrophys. 466, 743 (2007) The transiting planet OGLE-TR-132b revisited with new spectroscopy and deconvolution photometry.

J/A+A/466/763/  energies.dat  (22Kb)

Witthoeft M.C., Del Zanna G., Badnell N.R. Astron. Astrophys. 466, 763 (2007) Atomic Data from the IRON Project. LXIII. Electron-impact excitation of Fe19+ up to n=4.

J/A+A/466/771/  ionize.awk table29.dat  (156Kb)

Dere K.P. Astron. Astrophys. 466, 771 (2007) Ionization rate coefficients for the elements hydrogen through zinc.

J/A+A/466/823/  table3.dat  (22Kb)

Georgantopoulos I., Georgakakis A., Akylas A. Astron. Astrophys., 466, 823-830 (2007) Chandra and Spitzer observations of CDFS X-ray obscured QSOs.

J/A+A/466/905/  table.dat  (22Kb)

Fathi K., Beckman J.E., Zurita A., Relano M., Knapen J.H., Daigle O., Hernandez O., Carignan C. Astron. Astrophys. 466, 905 (2007) Evolution of structure in late-type spiral galaxies. I: Ionized gas kinematics in NGC 628.

J/A+A/466/917/  table1.dat table2.dat  (17Kb)

Caballero J.A. Astron. Astrophys. 466, 917 (2007)

The brightest stars of the sigma Orionis cluster.

J/A+A/466/931/  table1.dat  (4.6 Mb)

Carraro G., de la Fuente Marcos R., Villanova S., Moni Bidin C., de la Fuente Marcos C., Baumgardt H., Solivella G. Astron. Astrophys. 466, 931 (2007) Observational templates of star cluster disruption. The stellar group NGC 1901 in front of the Large Magellanic Cloud.

J/A+A/466/949/  tablea1.dat tablea2.dat tablea3.dat  (18Kb)

Nissen H.D., Gustafsson M., Lemaire J.L., Clenet Y., Rouan D., Field D. Astron. Astrophys., 466, 949-968 (2007) Observations of spatial and velocity structure in the Orion molecular cloud.

J/A+A/466/989/  fig/* list.dat  (14.2 Mb)

Beuther H., Walsh A.J., Throwirth S., Zhang Q., Hunter T.R., Megeath S.T., Menten K.M. Astron. Astrophys. 466, 989 (2007) Hot ammonia from NGC 6334I and I(N).

J/A+A/466/1031/ dao b.dat mdm b.dat mdm bg38.dat table1.dat ultrac g.dat ultrac i.dat ultrac u.dat uvic cle.dat uvic r.dat uvic v.dat (3.2 Mb) van den Besselaar E.J.M., Greimel R., Morales-Rueda L., Nelemans G., Thorstensen R., Marsh T.R., Dhillon V.S., Robb R.M., Balam D.D., Guenther E.W., Kemp J., Augusteijn T., Groot P.J. Astron. Astrophys. 466, 1031 (2007) DE CVn: a bright, eclipsing red dwarf - white dwarf binary.

J/A+A/466/1065/  figs/* list.dat  (6.6 Mb)

Beuther H., Leurini S., Schilke P., Wyrowski F., Menten K.M., Zhang Q. Astron. Astrophys. 466, 1065 (2007) Interferometric multi-wavelength (sub)millimeter continuum study of the young high-mass protocluster IRAS 05358+3543.

J/A+A/466/1099/  table4.dat  (13Kb)

Ohnaka K., Driebe T., Weigelt G., Wittkowski M. Astron. Astrophys. 466, 1099 (2007) Temporal variations of the outer atmosphere and the dust shell of the carbon-rich Mira variable V Oph probed with VLTI/MIDI.

J/A+A/466/1205/  images/* table1.dat  (632.1 Mb)

Stickel, M., Krause O., Klaas U., Lemke D. Astron. Astrophys. 466, 1205 (2007) The ISOPHOT 170micron Serendipity Survey. IV. The far-infrared sky atlas.

J/A+A/467/37/  table1.dat  (15Kb)

Barrena R., Boschin W., Girardi M., Spolaor M. Astron. Astrophys. 467, 37 (2007) Internal dynamics of the radio halo cluster A773.

J/A+A/467/73/  tablea1.dat  (24Kb)

Tajer M., Polletta M., Chiappetti L., Maraschi L., Trinchieri G., Maccagni D., Andreon S., Garcet O., Surdej J., Pierre M., Le Fevre O., Franceschini A., Lonsdale C.J., Surace J.A., Shupe D.L., Fang F., Rowan-Robinson M., Smith H.E., Tresse L. Astron. Astrophys., 467, 73-91 (2007) Obscured and unobscured AGN populations in a hard-X-ray selected sample of the XMDS survey.

J/A+A/467/229/  table2a.dat table2b.dat table3a.dat table3b.dat table5a.dat table5b.dat  (99Kb)

Panov K., Dimitrov D. Astron. Astrophys. 467, 229 (2007) Long-term photometric study of FK Comae and HD 199178.

J/A+A/467/565/  table1.dat table2.dat table3.dat table4.dat table5.dat table6.dat  (39Kb)

Berta S., Lonsdale C.J., Siana B., Farrah D., Smith H.E., Polletta M.C., Franceschini A., Fritz J., Perez-Fournon I., Rowan-Robinson M., Shupe D., Surace J. Astron. Astrophys., 467, 565-584 (2007) Keck spectroscopy of z=1-3 ULIRGs from the Spitzer SWIRE survey.

J/A+A/467/585/  table1.dat types.dat  (190Kb)

Bodaghee A., Courvoisier T.J.-L., Rodriguez J., Beckmann V., Produit N., Hannikainen D., Kuulkers E., Willis D.R., Wendt G. Astron. Astrophys., 467, 585-596 (2007) A description of sources detected by INTEGRAL during the first 4 years of observations.

J/A+A/467/665/  table1.dat table4.dat table5.dat  (15Kb)

Feltzing S., Fohlman M., Bensby T. Astron. Astrophys. 467, 665 (2007) Manganese trends in a sample of thin and thick disk stars. The origin of Mn.

J/A+A/467/785/  table1.dat table2.dat  (50Kb)

Norton A.J., Wheatley P.J., West R.G., Haswell C.A., Street R.A., Collier Cameron A., Christian D.J., Clarkson W.I., Enoch B., Gallaway M., Hellier C., Horne K., Irwin J., Kane S.R., Lister T.A., Nicholas J.P., Parley N., Pollacco D., Ryans R., Skillen I., Wilson D.M. Astron. Astrophys. 467, 785 (2007) New periodic variable stars coincident with ROSAT sources discovered using SuperWASP.

J/A+A/467/1003/  table4.dat  (22Kb)

Chies-Santos A.L., Santiago B.X., Pastoriza M.G. Astron. Astrophys. 467, 1003 (2007) High resolution imaging of the early type galaxy NGC 1380: an insight into the nature of extended star clusters

J/A+A/467/1025/  table2.dat table3.dat  (693Kb)

Gullieuszik M., Held E.V., Rizzi L., Saviane I., Momany Y., Ortolani S. Astron. Astrophys. 467, 1025 (2007) Near-infrared observations of the Fornax dwarf galaxy. I. The red giant branch.

J/A+A/467/1215/  table2.dat  (797Kb)

Southworth J., Bruntt H., Buzasi D.L. Astron. Astrophys. 467, 1215 (2007) Eclipsing binaries observed with the WIRE satellite. II. beta Aurigae and non-linear limb darkening in light curves.

J/A+A/467/1389/  all/* grid.dat z04.dat z06.dat z10.dat  (160.8 Mb)

Claret A. Astron. Astrophys. 467, 1389 (2007) New grids of stellar models including tidal-evolution constants up to carbon burning. IV. From 0.8 to 125 M_{⊙}: high metallicities (Z=0.04-0.10).

J/A+A/467/1397/  catalog.dat  (91Kb)

Delgado A.J., Alfaro E.J., Yun J.L. Astron. Astrophys. 467, 1397 (2007) Pre-main sequence stars in open clusters. I. The DAY-I Catalogue.

J/A+A/468/L13/  table2.dat  (18Kb)

Ortiz J.L., Santos-Sanz P., Gutierrez P.J., Duffard R., Aceituno F. Astron. Astrophys. 468, L13 (2007) Short term rotational variability in the large TNO 2005FY9.

J/A+A/468/49/  table2.dat  (31Kb)

Voss R., Gilfanov M. Astron. Astrophys., 468, 49-59 (2007) A study of the population of LMXBs in the bulge of M31.

J/A+A/468/151/  table2.dat  (31Kb)

Piskunov A.E., Schilbach E., Kharchenko N.V., Roeser S., Scholz R.-D. Astron. Astrophys. 468, 151 (2007) Towards absolute scales for the radii and masses of open clusters.

J/A+A/468/163/  refs.dat table7.dat table8.dat  (643Kb)

Deacon N.R., Hambly N.C. Astron. Astrophys. 468, 163 (2007) Southern infrared proper motion survey. II. A sample of low mass stars with mu 0.1”/yr.

J/A+A/468/205 See VI/125/

J/A+A/468/353/  table14.dat  (453Kb)

Guedel M., Briggs K.R., Arzner K., Audard M., Bouvier J., Feigelson E.D., Franciosini E., Glauser A., Grosso N., Micela G., Monin J.-L., Montmerle T., Padgett D.L., Palla F., Pillitteri I., Rebull L., Scelsi L., Silva B., Skinner S.L., Stelzer B., Telleschi A. Astron. Astrophys. 468, 353 (2007) The XMM-Newton Extended Survey of the Taurus Molecular Cloud.

J/A+A/468/379/  catalog.dat  (265Kb)

Audard M., Briggs K.R., Grosso N., Guedel M., Scelsi L., Bouvier J., Telleschi A. Astron. Astrophys. 468, 379 (2007)

The XMM-Newton Optical Monitor Survey of the Taurus Molecular Cloud.

J/A+A/468/637/  spm.dat table1.dat xl.dat zn.dat  (24.0 Mb)

Costa J.E.S., Michel E., Pena J., Creevey O., Li Z.P., Chevreton M., Belmonte J.A., Alvarez M., Fox Machado L., Parrao L.,F. Perez Hernandez F. Fernandez A., Fremy J.R., Pau S., Alonso R. Astron. Astrophys. 468, 637 (2007) Pulsational frequencies of the eclipsing delta-Scuti star HD 172189. Results of the STEPHI XIII campaign.

J/A+A/468/663/  table2.dat table3.dat table4.dat table5.dat table6.dat  (1.0 Mb)

Takeda Y., Kawanomoto S., Honda S., Ando H., Sakurai T. Astron. Astrophys. 468, 663 (2007) Behaviors of Li abundances in solar-analog stars: Evidence for line-width dependence.

J/A+A/468/807/  table2.dat  (20Kb)

Mannucci F., Bonnoli G., Maiolino R., Zappacosta L., Pedani M. Astron. Astrophys. 468, 807 (2007) A galaxy overdensity at z=0.401 associated with an X-ray emitting structure of warm-hot intergalactic medium.

J/A+A/468/937/  table6.dat  (11Kb)

Huertas-Company M., Rouan D., Soucail G., Le Fevre O., Tasca L., Contini T. Astron. Astrophys., 468, 937-950 (2007) Morphological evolution of z 1 galaxies from deep K-band AO imaging in the COSMOS deep field.

J/A+A/468/1001/  /20178 610.fits /20178 1280.fits table2.dat  (2.7 Mb)

Tej A., Ghosh S.K., Kulkarni V.K., Ojha D.K., Verma R.P., Vig S. Astron. Astrophys. 468, 1001 (2007) A multiwavelength study of the ultracompact HII region associated with IRAS 20178+4046.

J/A+A/468/1009/  refs.dat table4.dat table5.dat  (28Kb)

Hatchell J., Fuller G.A., Richer J.S., Harries T.J., Ladd E.F. Astron. Astrophys., 468, 1009-1024 (2007) Star formation in Perseus. II. SEDs, classification, and lifetimes.

J/A+A/468/1039/  phot.dat photmid.dat  (20Kb)

Broeg C. , Schmidt T. , Guenther E., Gaedke A., Bedalov A., Neuhauser R., Walter F.M. Astron. Astrophys. 468, 1039 (2007) Rotational period of GQ Lup.

J/A+A/468/1115/  table1.dat  (482Kb)

Lovis C., Pepe F. Astron. Astrophys. 468, 1115 (2007) A new list of thorium and argon spectral lines in the visible.

J/A+A/469/27/  table1.dat table23.dat  (900Kb)

Carrera F.J., Ebrer J., Mateos S., Ceballos M.T., Corral A., Barcons X., Page M.J., Rosen S.R., Watson M.G., Tedds J.A., Della Ceca R., Maccacaro T., Brunner H., Freyberg M., Lamer G., Bauer F.E., Ueda Y. Astron. Astrophys. 469, 27 (2007) The XMM-Newton serendipitous survey. III. The AXIS X-ray source counts and angular clustering.

J/A+A/469/173/  fits/* images.dat  (13.5 Mb)

Hily-Blant P., Falgarone E. Astron. Astrophys. 469, 173 (2007) Dissipative structures of diffuse molecular gas. II. The translucent environment of a dense core.

J/A+A/469/207/  fig1/* list.dat  (908Kb)

Comito C., Schilke P., Endesfelder U., Jimenez-Serra I., Martin-Pintado J. Astron. Astrophys. 469, 207 (2007) High-resolution study of a star-forming cluster in the Cep-A HW2 region.

J/A+A/469/285/  table5.dat  (46Kb)

Vaz L.P.R., Andersen J., Claret A. Astron. Astrophys. 469, 285 (2007) Absolute dimensions of eclipsing binaries. XXV. U Ophiuchi and the evolution and composition of 5 M_{⊙} stars.

J/A+A/469/309/  table2.dat  (21Kb)

Cincunegui C., Diaz R.F., Mauas P.J.D. Astron. Astrophys., 469, 309-317 (2007) Hαand the CaII H and K lines as activity proxies for late-type stars.

J/A+A/469/451/  table2.dat  (11Kb)

Torniainen I., Tornikoski M., Lahteenmaki A., Aller M.F., Aller H.D., Mingaliev M.G. Astron. Astrophys. 469, 451 (2007)

Radio continuum spectra of gigahertz-peaked spectrum galaxies.

J/A+A/469/459/  table3.dat  (19Kb)

Testor G., Lemaire J.L., Kristensen L.E., Field D., Diana S. Astron. Astrophys., 469, 459-469 (2007) VLT/NACO near-infrared imaging and spectroscopy of N159-5 in the LMC HII complex N159.

J/A+A/469/503/  tablea1.dat  (69Kb)

Dolcini A., Farfanelli F., Ciprini S., Treves A., Covino S., Tosti G., Pian E., Sbarufatti B., Molinari E., Chincarini G., Zerbi F.M., Conconi P., Nicastro L., Palazzi E., Testa V., Vitali F., Antonelli L.A., Danziger J., Tagliaferri G., Meurs E., Vergani S., Fernandez-Soto A., Distefano E., Cutispoto G., D’Alessio F. Astron. Astrophys. 469, 503 (2007) REM near-IR and optical multiband observations of PKS 2155-304 in 2005.

J/A+A/469/529/  m75phot.dat  (142Kb)

Kravtsov, V., Alcaino, G., Marconi, G., Alvarado F. Astron. Astrophys. 469, 529 (2007) Multi-color photometry of the Galactic globular cluster M75 = NGC 6864. A new sensitive metallicity indicator and the position of the horizontal branch in UV.

J/A+A/469/575/ fig1sig1.fits fig1sig2.fits fig2sig1.fits fig2sig2.fits fig3sig1.fits fig3sig2.fits fig4sig1.fits fig4sig2.fits fig4sig3.fits list.dat (64Kb) Berne O., Joblin C., Deville Y., Smith J.D., Rapacioli M., Bernard J.P., Thomas J., Reach W., Abergel A. Astron. Astrophys. 469, 575 (2007) Analysis of the emission of very small dust particles from Spitzer spectro-imagery data using blind signal separation methods.

J/A+A/469/607/  table2.dat  (25Kb)

Weltevrede P., Stappers B.W., Edwards R.T. Astron. Astrophys. 469, 607 (2007) The subpulse modulation properties of pulsars at 92 cm and the frequency dependence of subpulse modulation.

J/A+A/469/713/  fig/* table12.dat table35.dat  (15.4 Mb)

Medhi B.J., Messina S., Parihar P.S., Pagano I., Muneer S., Duorah K. Astron. Astrophys. 469, 713 (2007) Results from a spectroscopic survey in the CoRoT fields. I. Search for chromospherically active stars.

J/A+A/469/783/  table2.dat  (16Kb)

Sousa S.G., Santos N.C., Israelian G., Mayor M., Monteiro M.J.P.F.G. Astron. Astrophys. 469, 783 (2007) A new code for automatic determination of equivalent widths: Automatic Routine for line Equivalent widths in stellar Spectra (ARES).

J/A+A/469/799/  candid.dat catalog.dat disqual.dat  (257Kb)

Szczerba R., Siodmiak N., Stasinska G., Borkowski J. Astron. Astrophys., 469, 799-806 (2007) An evolutionary catalogue of galactic post-AGB and related objects.

J/A+A/469/807/  lmxb.dat lmxbnote.dat refs.dat  (226Kb)

Liu Q.Z., van Paradijs J., van den Heuvel E.P.J. Astron. Astrophys. 469, 807 (2007) Catalogue of low-mass X-ray binaries in the Galaxy (4th edition).

J/A+A/469/861/  table1.dat  (13Kb)

Barrena R., Boschin W., Girardi M., Spolaor M. Astron. Astrophys., 469, 861-872 (2007) The dynamical status of the galaxy cluster Abell 115.

J/A+A/469/1033/ table4.dat table5.dat table6.dat table7.dat table8.dat table9.dat table10.dat table11.dat table12.dat table13.dat table14.dat table15.dat (116Kb) Sodor A., Szaidl B., Jurcsik J. Astron. Astrophys. 469, 1033 (2007) The Blazhko behaviour of RR Geminorum. II. Long-term photometric results.

J/A+A/469/1069/  rv.dat  (12Kb)

Butkovskaya V., Plachinda S. Astron. Astrophys. 469, 1069 (2007) A study of the beta Cephei star gamma Pegasi: binarity, magnetic field, rotation and pulsations.

J/A+A/469/1203/  avalues.dat energyle.dat table1.dat table2.dat upsilon.dat  (191Kb)

Melendez M., Bautista M.A., Badnell N.R.

Astron. Astrophys. 469, 1203 (2007) Atomic data from the IRON project. LXIV. Radiative transition rates and collision strengths for Ca II.

J/A+A/469/1211/  table2.dat  (15Kb)

Trevese D., Vagnetti F., Puccetti S., Fiore F., Tomei M., Bershady M.A. Astron. Astrophys. 469, 1211 (2007) An X-ray Survey in SA 57 with XMM-Newton.

J/A+A/469/1221/  ac/* merid/* misc/* sogs/* table1.dat table2.dat table3.dat  (71.7 Mb)

Fresneau A., Vaughan A.E., Argyle R.W. Astron. Astrophys. 469, 1221 (2007) Sydney observatory Galactic survey. Potential for detection of extreme disk-crossing stars.

J/A+A/470/39/  tablea1.dat  (15Kb)

Ramella M., Biviano A., Pisani A., Varela J., Bettoni D., Couch W.J., D’onofrio M., Dressler A., Fasano G., Kjaergaard P., Moles M., Pignatelli E., Poggianti B.M. Astron. Astrophys., 470, 39-51 (2007) Substructures in WINGS clusters.

J/A+A/470/161/  table2.dat table3.dat  (15Kb)

Russeil D., Adami C., Georgelin Y.M. Astron. Astrophys., 470, 161-171 (2007) Revised distances of Northern HII regions.

J/A+A/470/191/  tablea1.dat tablea2.dat tablea3.dat tablea4.dat  (35Kb)

Wittkowski M., Boboltz D.A., Ohnaka K., Driebe T., Scholz M. Astron. Astrophys. 470, 191 (2007) The Mira variable S Orionis: relationships between the photosphere, molecular layer, dust shell, and SiO maser shell at 4 epochs.

J/A+A/470/231/  H.fits I.fits J.fits K.fits L.fits table1.dat  (4.9 Mb)

Yun J.L., Torrelles J.M., Santos N.C. Astron. Astrophys. 470, 231 (2007) A diamond-ring star: the unusual morphology of a young (multiple?) object.

J/A+A/470/281/  refs.dat table7.dat  (20Kb)

Spezzi L., Alcala J.M., Frasca A., Covino E., Gandolfi D. Astron. Astrophys. 470, 281 (2007) A WFI survey in the Chamaeleon II dark cloud.

J/A+A/470/317/ figurea1.png figurea2.png tablea1.dat tablea2.dat tablea3.dat tablea4.dat tablea5.dat (8.2 Mb) Rauch T., Ziegler M., Werner K., Kruk J.W., Oliveira C.M., Vande Putte D., Migniani R.P., Kerber F. Astron. Astrophys. 470, 317 (2007) High-resolution FUSE and HST ultraviolet spectroscopy of the white dwarf central star of Sh 2-216.

J/A+A/470/505/  table3.dat  (3.8 Mb)

Verley S., Odewahn S. C., Verdes-Montenegro L., Leon S., Combes F., Sulentic J., Bergond G., Espada D., Garcia E., Lisenfeld U., Sabater J. Astron. Astrophys. 470, 505 (2007) The AMIGA sample of isolated galaxies. IV. A catalogue of neighbours around isolated galaxies.

J/A+A/470/585/  table3.dat table4.dat table7.dat  (471Kb)

Balaguer-Nunez L., Galadi-Enriquez D., Jordi C., Astron. Astrophys. 470, 585 (2007) uvby-H_{β} CCD photometry and membership segregation of the open cluster NGC 2682 (M 67).

J/A+A/470/605/  HLTau Ha.fit HLTau SII.fit images.dat  (30.6 Mb)

Movsessian T.A., Magakian T.Yu., Bally J., Smith M.D., Moiseev A.V., Dodonov S.N. Astron. Astrophys. 470, 605 (2007) Herbig-Haro jets in 3D: the HL/XZ Tauri region.

J/A+A/470/787/  table3.dat table4.dat table6.dat  (23Kb)

Piranomonte S., Perri M., Giommi P., Landt H., Padovani P. Astron. Astrophys., 470, 787-809 (2007) The sedentary survey of extreme high-energy peaked BL Lacs. III. Results from optical spectroscopy.

J/A+A/470/821/  notes.dat table1.dat table2.dat table3.dat table4.dat  (37Kb)

Dietrich J.P., Erben T., Lamer G., Schneider P., Schwope A., Hartlap J., Maturi M. Astron. Astrophys. 470, 821 (2007) BLOX: The Bonn Lensing, Optical, and X-ray selected galaxy clusters - I. Cluster catalog construction.

J/A+A/470/957/  fits/* images.dat  (4.3 Mb)

Murakawa K., Nakashima J., Ohnaka K., Deguchi S. Astron. Astrophys. 470, 957 (2007) Near-infrared polarimetric study of the bipolar nebula IRAS 19312+1950.

J/A+A/470/1009/  table.dat  (51Kb)

Carrier F., Eggenberger P., Leyder J.C., Debernardi Y., Royer F. Astron. Astrophys. 470, 1009 (2007) A search for solar-like oscillations in the Am star HD 209625.

J/A+A/471/L51/  phot.dat  (21Kb)

Gillon M., Demory B.-O., Barman T., Bonfils X., Mazeh T., Pont F., Udry S., Mayor M., Queloz D. Astron. Astrophys. 471, L51 (2007) Accurate Spitzer infrared radius measurement for the hot Neptune GJ 436b.

J/A+A/471/17/  table1.dat  (12Kb)

Aguerri J.A.L., Sanchez-Janssen R., Munoz-Tunon C. Astron. Astrophys. 470, 17 (2007) A study of catalogued nearby clusters in the SDSS-DR4. I. Cluster global properties.

J/A+A/471/103/  smcsimbad.fit  (447Kb)

Bot C., Boulanger F., Rubio M., Rantakyro F. Astron. Astrophys. 471, 103 (2007) Millimeter dust continuum emission revealing the true mass of giant molecular clouds in the Small Magellanic Cloud

J/A+A/471/165/  table1.dat table2.dat  (17Kb)

Orsatti A.M., Feinstein C., Vega E.I., Vergne M.M. Astron. Astrophys., 471, 165-172 (2007) Revisiting the region of the open cluster NGC 5606.

J/A+A/471/499/  lmscand.dat table3.dat  (98Kb)

Moraux E., Bouvier J., Stauffer J.R., Barrado y Navascues D., Cuillandre J.C. Astron. Astrophys. 471, 499 (2007) The lower mass function of the young open cluster Blanco 1: from 30M_{Jup} to 3M_{⊙}.

J/A+A/471/515/  lc/* table.dat  (842Kb)

De Marchi F., Poretti E., Montalto M., Piotto G., Desidera S., Bedin L.R., Claudi R., Arellano Ferro A., Bruntt H., Stetson P.B. Astron. Astrophys. 471, 515 (2007) Variable stars in the open cluster NGC 6791 and its surrounding field.

J/A+A/471/573/  table1.dat  (5Kb)

Yun J.L., Lopez-Sepulcre A., Torrelles J.M. Astron. Astrophys. 471, 573 (2007) A new young stellar cluster embedded in a molecular cloud in the far outer Galaxy.

J/A+A/471/625/ stars.dat tablea1.dat tablea2.dat tableb1.dat tablec1.dat tablec2.dat tablec3.dat tablec4.dat (188Kb) Trundle C., Dufton P.L., Hunter I., Evans C.J., Lennon D.J., Smartt S.J., Ryans R.S.I. Astron. Astrophys., 471, 625-643 (2007) The VLT-FLAMES survey of massive stars: Evolution of surface N abundances and effective temperature scales in the Galaxy and Magellanic Clouds.

J/A+A/471/795/  comp.dat dev.dat  (1.4 Mb)

Li Z., Han Z. Astron. Astrophys. 471, 795 (2007) How young stellar populations affect the ages and metallicities of galaxies.

J/A+A/471/813/  table3.dat  (15Kb)

Roman-Lopes A. Astron. Astrophys., 471, 813-824 (2007) Near-infrared study of southern massive star formation regions. The case of IRAS 16571-4029 source.

J/A+A/471/1069/ lea-406.for lea-406a.for lea-406b.for table6.dat table7.dat table8.dat table9.dat table10.dat table11.dat (4.2 Mb) Kudryavtsev S.M. Astron. Astrophys. 471, 1069 (2007) Long-term harmonic development of lunar ephemeris.

J/A+A/471/1105/  table3.dat table4.dat table5.dat  (237Kb)

Tasse C., Roettgering H.J.A., Best P.N., Cohen A.S., Pierre M., Wilman R. Astron. Astrophys. 471, 1105 (2007) GMRT observations of the XMM Large Scale Structure Survey field.

J/A+A/472/101/  object.dat table2.dat table3.dat table7.dat  (82Kb)

Idiart T.P., Maciel W.J., Costa R.D.D. Astron. Astrophys. 472, 101 (2007) Chemical evolution of the Small Magellanic Cloud based on planetary nebulae.

J/A+A/472/111/  table1.dat table2.dat  (11Kb)

Mieske S., Hilker M., Jordan A., Infante L., Kissler-Patig M. Astron. Astrophys., 472, 111-119 (2007) A search for ultra-compact dwarf galaxies in the Centaurus galaxy cluster.

J/A+A/472/121/  tables.dat  (95Kb)

Verley S., Leon S., Verdes-Montenegro L., Combes F., Sabater J., Sulentic J., Bergond G., Espada D., Garcia E., Lisenfeld U., Odewhan S.C. Astron. Astrophys. 472, 121 (2007) The AMIGA sample of isolated galaxies. V. Quantification of the isolation.

J/A+A/472/187/  fits/* images.dat objects.dat  (47.1 Mb)

Hatchell J., Fuller G.A., Richer J.S. Astron. Astrophys. 472, 187 (2007) Star Formation in Perseus: III. Outflows.

J/A+A/472/241/  notes.dat table1.dat table2.dat table3.dat  (17Kb)

Gerbaldi M., Faraggiana R., Caffau E. Astron. Astrophys. 472, 241 (2007) UV Flux distributions of gamma Dor stars.

J/A+A/472/519/  fits/* object.dat  (5.5 Mb)

Andre P., Belloche A., Motte F., Peretto N. Astron. Astrophys. 472, 519 (2007) The initial conditions of star formation in the Ophiuchus main cloud: Kinematics of the protocluster condensations.

J/A+A/472/557/  table3.dat  (21Kb)

Marti J., Paredes J.M., Ishwara Chandra C.H., Bosch-Ramon V. Astron. Astrophys., 472, 557-564 (2007) Deep radio images of the HEGRA and Whipple TeV sources in the Cygnus OB2 region.

J/A+A/472/565/  tablea1.dat tablea2.dat  (19Kb)

Gutierrez-Soto J., Fabregat J., Suso J., Suarez J.C., Moya A., Garrido R., Hubert A.-M., Floquet M., Neiner C., Fremat Y. Astron. Astrophys. 472, 565 (2007) Multiperiodic pulsations in the Be stars NW Ser and V1446 Aql.

J/A+A/472/577/  table1.dat  (22Kb)

Martayan C., Floquet M., Hubert A.M., Gutierrez-Soto J., Fabregat J., Neiner C., Mekkas M. Astron. Astrophys., 472, 577-586 (2007) Be stars and binaries in the field of the SMC open cluster NGC 330 with VLT-FLAMES.

J/A+A/472/699/  catalog.dat  (92Kb)

Turriziani S., Cavazzuti E., Giommi P. Astron. Astrophys. 472, 699 (2007) ROXA: a new multi-frequency large sample of blazars selected with SDSS and 2dF optical spectroscopy.

J/A+A/472/705/  table3.dat  (49Kb)

Verrecchia F., in ’t Zand J.J.M., Giommi P., Santolamazza P., Granata S., Schuurmans J.J., Antonelli A. Astron. Astrophys. 472, 705 (2007) The BeppoSAX WFC X-ray source catalogue

J/A+A/472/763/  fig/* sources.dat table2.dat  (9.8 Mb)

Britzen S., Vermeulen R.C., Taylor G.B., Campbell R.M., Pearson T.J., Readhead A.C.S., Xu W., Browne I.W.A., Henstock D.R., Wilkinson P. Astron. Astrophys. 472, 763 (2007) A multi-epoch VLBI survey of the kinematics of CJF sources Part I: Model-fit parameters and maps.

J/A+A/472/785/  m33-3.6cmI.fits  (50Kb)

Tabatabaei F.S., Krause M., Beck R. Astron. Astrophys. 472, 785 (2007)

High-resolution radio continuum survey of M33. I. The radio maps.

J/A+A/472/805/  table12.dat  (25Kb)

Richards A.M.S., Muxlow T.W.B., Beswick R., Allen M.G., Benson K., Dickson R.C., Garrett M.A., Garrington S.T., Gonzalez-Solarez E., Harrison P.A., Holloway A.J., Kettenis M.M., Laing R.A., Richards E.A., Thrall H., van Langevelde H.J., Walton N.A., Wilkinson P.N., Winstanley N. Astron. Astrophys. 472, 805 (2007) Using VO tools to investigate distant radio starbursts hosting obscured AGN in the HDF(N) region.

J/A+A/472/1029/ oh2olev.dat oh2ooh2.dat oh2oph2.dat ph2olev.dat ph2ooh2.dat ph2oph2.dat (329Kb) Faure A., Crimier N., Ceccarelli C., Valiron P., Wiesenfeld L., Dubernet L. Astron. Astrophys. 472, 1029 (2007) Quasi-classical rate coefficient calculations for the rotational (de)excitation of H2O by H2.

J/A+A/472/1055/  table1.dat table2.dat table3.dat  (46Kb)

Sinachopoulos D., Gavras P., Dionatos O., Ducourant C., Medupe T. Astron. Astrophys. 472, 1055 (2007) CCD astrometry and components instrumental magnitude difference of 432 Hipparcos wide visual double stars.

J/A+A/473/143/  table1.dat table4.dat  (17Kb)

Demers S., Battinelli P. Astron. Astrophys., 473, 143-148 (2007) C stars as kinematic probes of the Milky Way disk from 9 to 15 kpc.

J/A+A/473/149/  table2.dat  (19Kb)

Comeron F., Schneider N. Astron. Astrophys., 473, 149-162 (2007) A close look at the heart of RCW 108.

J/A+A/473/177/  fits/* list.dat  (22Kb)

Wirstrom E.S., Bergman P., Hjalmarson A., Nummelin A. Astron. Astrophys. 473, 177 (2007) A search for pre-biotic molecules in hot cores.

J/A+A/473/347/  table1.dat table2.dat  (14Kb)

Vetere L., Soffitta P., Massaro E., Giommi P., Costa E. Astron. Astrophys., 473, 347-349 (2007) The complete catalogue of GRBs observed by the wide field cameras on board BeppoSAX.

J/A+A/473/369/  refs.dat table3.dat  (24Kb)

Haas M., Siebenmorgen R., Pantin E., Horst H., Smette A., Kaufl H.-U., Lagage P.-O., Chini R. Astron. Astrophys. 473, 369 (2007) VISIR/VLT mid-infrared imaging of Seyfert nuclei: nuclear dust emission and the Seyfert-2 dichotomy.

J/A+A/473/399/  table1.dat table8.dat table9.dat  (15Kb)

Pompei E., Dahlem M., Iovino A. Astron. Astrophys., 473, 399-409 (2007) Optical and radio survey of southern compact groups of galaxies. I. Pilot study of six groups.

J/A+A/473/429/  table5.dat  (150Kb)

Federici L., Bellazzini M., Galleti S., Fusi Pecci F., Buzzoni A., Parmeggiani G. Astron. Astrophys. 473, 429 (2007) The extended structure of the remote cluster B514 in M31. Detection of extra-tidal stars.

J/A+A/473/523/  cenx3/* list.dat lmcx4/* smcx1/*  (143Kb)

van der Meer A., Kaper L., van Kerkwijk M.H., Heemskerk M.H.M., van den Heuvel E.P.J. Astron. Astrophys. 473, 523 (2007) Determination of the mass of the neutron star in SMC X-1, LMC X-4 and Cen X-3 with VLT/UVES.

J/A+A/473/791/  position.dat table1.dat  (16Kb)

Fox A.J., Ledoux C., Petitjean P., Srianand R. Astron. Astrophys. 473, 791 (2007) C IV absorption in damped and sub-damped Lyman-alpha systems. Correlations with metallicity and implications for galactic winds at z= 2-3.

J/A+A/473/829/  catalog.dat cfa.dat cfapos.dat  (56Kb)

Mermilliod J.-C., Andersen J., Latham D., Mayor M. Astron. Astrophys. 473, 829 (2007) Red giants in open clusters. XIII. Orbital elements of 156 spectroscopic binaries.

J/A+A/473/995/  table1.dat table3.dat table4.dat table5.dat table6.dat tables.ps tables.tex  (2.0 Mb)

Aggarwal K.M., Keenan F.P., Msezane A.Z.

Astron. Astrophys. 473, 995 (2007) Energy levels and radiative rates for transitions in Co XI.

J/A+A/474/121/  phot.dat table1.dat  (73Kb)

Borissova J., Ivanov V.D., Stephens A.W., Catelan M., Minniti M., Prieto G. Astron. Astrophys. 474, 121 (2007) A deep near-infrared view of the Galactic globular cluster 2MASS GC02.

J/A+A/474/293/  table1.dat table2.dat table3.dat  (11Kb)

Bonfils X., Mayor M., Delfosse X., Forveille T., Gillon M., Perrier C., Udry S., Bouchy F., Lovis C., Pepe F., Queloz D., Santos N.C., Bertaux J.-L. Astron. Astrophys. 474, 293 (2007) The HARPS search for southern extra-solar planets - X. A msini=11M_{⊕} planet around the nearby spotted M dwarf GJ 674.

J/A+A/474/473/  table4.dat  (22Kb)

Garcet O., Gandhi P., Gosset E., Sprimont P.G., Surdej J., Borkowski V., Tajer M., Pacaud F., Pierre M., Chiappetti L., Maccagni D., Page M.J., Carrera F.J., Tedds J.A., Mateos S., Krumpe M., Contini T., Corral A., Ebrero J., Gavignaud I., Schwope A., Le Fevre O., Polletta M., Rosen S., Lonsdale C., Watson M., Borczyk W., Vaisanen P. Astron. Astrophys., 474, 473-489 (2007) The XMM large scale structure survey: optical vs. X-ray classifications of active galactic nuclei and the unified scheme.

J/A+A/474/495/  table1.dat  (18Kb)

Albacete Colombo J.F., Caramazza M., Flaccomio E., Micela G., Sciortino S. Astron. Astrophys., 474, 495-504 (2007) X-ray flaring from the young stars in Cygnus OB2.

J/A+A/474/763/  table1.dat  (25Kb)

Jablonka P., Gorgas J., Goudfrooij P. Astron. Astrophys., 474, 763-776 (2007) Stellar population gradients in bulges along the Hubble sequence. II. Relations with galaxy properties.

J/A+A/474/783/  table2.dat  (723Kb)

Deng X.-F., He J.-Z., Jiang P., Luo C.-H., Wu P. Astron. Astrophys. 474, 783 (2007) The main galaxy groups from the SDSS Data Release 5.

J/A+A/474/793/  table1.dat  (6Kb)

Guidorzi C., Vergani S.D., Sazonov S., Covino S., Malesani D., Molkov S., Palazzi E., Romano P., Campana S., Chincarini G., Fugazza D., Moretti A., Tagliaferri G., Llorente A., Gorosabel J., Antonelli L.A., Capalbi M., Cusumano G., D’Avanzo P., Mangano V., Masetti N., Meurs E., Mineo T., Molinari E., Morris D.C., Nicastro L., Page K.L., Perri M., Sbarufatti B., Stratta G., Sunyaev R., Troja E., Zerbi F.M. Astron. Astrophys., 474, 793-805 (2007) GRB 070311: a direct link between the prompt emission and the afterglow.

J/A+A/474/883/  fits/* maps.dat  (55.2 Mb)

Birkmann S.M., Krause O., Hennemann M., Henning T., Steinacker J., Lemke D. Astron. Astrophys. 474, 883 (2007) A massive protostellar core with an infalling envelope.

J/A+A/474/891/  sp/* table3.dat  (13.1 Mb)

Urquhart J.S., Busfield A.L., Hoare M.G., Lumsden S.L., Oudmaijer R.D., Moore T.J.T., Gibb A.G., Purcell C.R., Burton M.G., Marechal L.J.L. Astron. Astrophys., 474, 891-901 (2007) The RMS survey. ^{13}CO observations of candidate massive YSOs in the southern Galactic plane.

J/A+A/474/911/  fits/* maps.dat  (60.2 Mb)

Palau A., Estalella R., Ho P.T.P., Beuther H., Beltran M. T. Astron. Astrophys. 474, 911 (2007) Unveiling the nature and interaction of the intermediate/high-mass YSOs in IRAS 20343+4129.

J/A+A/474/1093/  table1.dat  (25Kb)

Lowe S.R., Gawronski M.P., Wilkinson P.N., Kus A.J., Browne I.W.A., Pazderski E., Feiler R., Kettle D. Astron. Astrophys. 474, 1093 (2007) 30 GHz flux density measurements of the Caltech-Jodrell flat-spectrum sources with OCRA-p.

J/A+A/475/217/  refs.dat table1.dat table6.dat  (142Kb)

Quireza C., Rocha-Pinto H.J., Maciel W.J. Astron. Astrophys. 475, 217 (2007)

Bayesian posterior classification of planetary nebulae according to the Peimbert types.

J/A+A/475/281/  fit/* images.dat  (43.8 Mb)

Galfalk M., Olofsson G. Astron. Astrophys. 475, 281 (2007) Herbig-Haro flows in B335.

J/A+A/475/359/  table1.dat  (21Kb)

Griessmeier J.-M., Zarka P., Spreeuw H. Astron. Astrophys. 475, 359 (2007) Predicting low-frequency radio fluxes of known extrasolar planets.

J/A+A/475/393/  table1.dat table3.dat table4.dat table7.dat tables.tex  (5.5 Mb)

Aggarwal M., Keenan F.P. Astron. Astrophys. 475, 393 (2007) Energy levels, radiative rates and excitation rates for transitions in Ni XI.

J/A+A/475/539/  table4.dat  (22Kb)

Prisinzano L., Randich S. Astron. Astrophys. 475, 539 (2007) Lithium abundances in the old open cluster NGC 3960 from VLT/FLAMES observations.

J/A+A/475/549/  fits/* table1.dat  (67.7 Mb)

Benz A.O., Stauber P., Bourke T.L., van det Tak F.F.S., van Dishoeck E.F., Jorgensen J.K. Astron. Astrophys. 475, 549 (2007) Energetic radiation and the sulfur chemistry of protostellar envelopes: submillimeter interferometry of AFGL 2591.

J/A+A/475/559/  fits/* list.dat  (11.0 Mb)

Winters J.M., Le Bertre T., Pety J., Neri R. Astron. Astrophys. 475, 559 (2007) Mass loss from dusty, low outflow-velocity AGB stars. II. The multiple wind of EP Aquarii.

J/A+A/475/765/  table1.dat table3.dat  (35Kb)

Ramsbottom C.A., Hudson C.E., Norrington P.H., Scott M.P. Astron. Astrophys. 475, 765 (2007) Electron-impact excitation of Fe II. Collision strengths and effective collision strengths for low-lying finestructure forbidden transitions.

J/A+A/475/775/  table1.dat table2.dat  (43Kb)

Krivonos R., Revnivtsev M., Lutovinov A., Sazonov S., Churazov E., Sunyaev R. Astron. Astrophys. 475, 775 (2007) Catalog of hard X-ray sources detected in all-sky survey performed with INTEGRAL observatory.

J/A+A/475/791/  catalog.dat  (23Kb)

Seymour N., Rocca-Volmerange B., de Lapparent V. Astron. Astrophys. 475, 791 (2007) A 12 μm ISOCAM survey of the ESO-Sculptor field.

J/A+A/475/915/  fits/* list.dat  (103.4 Mb)

Brinch, C., Crapsi, A., Jorgensen, J., K., Hogerheijde, M., R., Hill, T. Astron. Astrophys. 475, 915 (2007) A deeply embedded young protoplanetary disk around L1489 IRS observed by the submillimeter array.

J/A+A/475/925/  figs/* list.dat  (560.5 Mb)

Leurini S., Beuther H., Schilke P., Wyrowski F., Menten K.M., Zhang Q. Astron. Astrophys. 475, 925 (2007) Multi-line (sub)millimetre observations of the high-mass proto cluster IRAS 05358+3543.

J/A+A/475/1003/  table4.dat  (31Kb)

Hekker S., Melendez J. Astron. Astrophys. 475, 1003 (2007) Precise radial velocities of giant stars. III. Spectroscopic stellar parameters.

J/A+A/475/1053/  stars.dat table3.dat  (27Kb)

Auriere M., Wade G.A., Silvester J., Lignieres F., Bagnulo S., Bale K., Dintrans B., Donati J.F., Folsom C.P., Gruberbauer M., Hui Bon Hoa A., Jeffers S., Johnson N., Landstreet J.D., Lebre A., Lueftinger T., Marsden S., Mouillet D., Naseri S., Paletou F., Petit P., Power J., Rincon F., Strasser S., Toque N. Astron. Astrophys. 475, 1053 (2007) Weak magnetic fields in Ap/Bp stars: Evidence for a dipole field lower limit and a tentative interpretation of the magnetic dichotomy.

J/A+A/475/1125/  caiihk.dat gbphot.dat sephot.dat  (41Kb)

Demory B.-O., Gillon M., Barman T., Bonfils X., Mayor M., Mazeh T., Queloz D., Udry S., Bouchy F., Delfosse X., Forveille T., Mallmann F., Pepe F., Perrier C. Astron. Astrophys. 475, 1125 (2007) Characterization of the hot Neptune GJ 436b with Spitzer and ground-based obs.

J/A+A/475/1159/  codes.dat defatts.dat gmclass.f gmreadme.txt init.dat list.dat  (942Kb)

Debosscher J., Sarro L.M., Aerts C., Cuypers J., Vandenbussche B., Garrido R., Solano E. Astron. Astrophys., 475, 1159-1183 (2007) Automated supervised classification of variable stars. I. Methodology.

J/A+A/476/59/  psc-fcss.dat xsc-fcss.dat  (31Kb)

Morris R.A.H, Phillipps S., Jones J.B., Drinkwater M.J., Gregg M.D, Couch W.J., Parker Q.A., Smith R.M. Astron. Astrophys. 476, 59 (2007) 2MASS Galaxies in the Fornax Cluster Spectroscopic Survey.

J/A+A/476/199/  naco.dat sofi.dat  (381Kb)

Ascenso J., Alves J., Vicente, S., Lago M.T.V.T. Astron. Astrophys. 476, 199 (2007) NTT and VLT diffraction limited imaging of Trumpler 14: revealing a massive core-halo cluster.

J/A+A/476/881/  table2a.dat table2b.dat table2c.dat  (62Kb)

Korhonen H., Berdyugina S.V., Hackman T., Ilyin I.V., Strassmeier K.G., Tuominen I. Astron. Astrophys. 476, 881 (2007) Study of FK Comae Berenices. V. Spot evolution and detection of surface differential rotation.

J/A+A/476/1243/  fits/* refs.dat table1.dat table2.dat  (6.8 Mb)

Motte F., Bontemps S., Schilke P., Schneider N., Menten K.M., Broguiere D. Astron. Astrophys. 476, 1243 (2007) The earliest phases of high-mass star formation: a 3 square degree millimeter continuum mapping of Cygnus X.

J/A+A/476/1347/  table1.dat  (22Kb)

Pont F., Gilliland R.L., Moutou C., Charbonneau D., Bouchy F., Brown T.M., Mayor M., Queloz D., Santos N., Udry S. Astron. Astrophys., 476, 1347-1355 (2007) Hubble Space Telescope time-series photometry of the planetary transit of HD 189733: no moon, no rings, starspots.

J/A+A/477/165/  table.dat  (85Kb)

Piskunov A.E., Schilbach E., Kharchenko N.V., Roeser S., Scholz R.-D. Astron. Astrophys. 477, 165 (2008) Tidal radii and masses of open clusters.

J/A+A/477/593/  table2.dat table3.dat  (212Kb)

Antokhin I.I., Rauw G., Vreux J.-M., van der Hucht K.A., Brown J.C. Astron. Astrophys. 477, 593 (2008) X-ray study of early type stars in the Carina OB1 association

J/A+A/477/877/  object.dat sp/*  (534Kb)

Pereira C.B., Marcolino W.L.F., Machado M., Araujo F.X. Astron. Astrophys. 477, 877 (2008) Spectroscopic analysis of two peculiar emission-line stars: RJHA 49 and SS73 21.

J/A+A/477/907/  lc1.dat stars.dat  (695Kb)

Pigulski A., Pojmanski G. Astron. Astrophys. 477, 907 (2008) Beta Cephei stars in the ASAS-3 data. I. Long-term variations of periods and amplitudes.

J/A+A/477/917/  lc2.dat stars.dat  (2.0 Mb)

Pigulski A., Pojmanski G. Astron. Astrophys. 477, 917 (2008) Beta Cephei stars in the ASAS-3 data. II. 103 new Beta Cephei stars and a discussion of low-frequency modes.

J/A+A/477/967/  DLSFIT demo.f90  (33Kb)

Bukvic S., Spasojevic Dj., Zigman V. Astron. Astrophys. 477, 967 (2008) Advanced fit technique for astrophysical spectra.

J/A+A/478/83/  table1.dat table1.fit  (170Kb)

Vanzella E., Cristiani S., Dickinson M., Giavalisco M., Kuntschner H., Haase J., Nonino M., Rosati P., Cesarsky C., Ferguson H.C., Fosbury R.A.E., Grazian A., Moustakas L.A., Rettura A., Popesso P., Renzini A., Stern D., GOODS Team Astron. Astrophys. 478, 83 (2008) The Great Observatories Origins Deep Survey. VLT/FORS2 Spectroscopy in the GOODS-South Field: part III.

J/A+A/478/197/ table1.dat (189Kb) Mosser B., Bouchy F., Martic M., Appourchaux T., Barban C., Berthomieu G., Garcia R.A., Lebrun J.C., Michel E., Provost J., Thevenin F., Turck-Chieze S. Astron. Astrophys. 478, 197 (2008) Asteroseismology of Procyon with SOPHIE.

J/A+A/478/321/ refs.dat table1.dat (27Kb) Borguet B., Hutsemekers D., Letawe G., Letawe Y., Magain P. Astron. Astrophys. 478, 321 (2008) Evidence of a Type1/Type2 dichotomy in the correlation between quasar optical polarization and host galaxy/extended emission position angles.

J/A+A/478/507/ table1.dat table2.dat (46Kb) Morales J.C., Ribas I., Jordi C. Astron. Astrophys. 478, 507 (2008) The effect of activity on stellar temperatures and radii.

J/A+A/478/659/ table4.dat table5.dat (1.7 Mb) Sabogal B.E., Mennickent R.E., Pietrzynski G., Garca J.A., Gieren W., Kolaczkowski Z. Astron. Astrophys. 478, 659 (2008) A catalogue of Be stars in the direction of the Galactic Bulge.

J/A+A/478/667/ tablea1.dat tablea2.dat tablea3.dat tablea4.dat tablea5.dat tablea6.dat tablea7.dat (93Kb) Caballero J.A. Astron. Astrophys. 478, 667 (2008) Stars and brown dwarfs in the sigma Orionis cluster: the Mayrit catalogue.

J/A+AS/1/189 See I/23/ J/A+AS/1/199 See II/21A/ J/A+AS/1/281 See VIII/36/ J/A+AS/2/69 See I/14/ J/A+AS/5/129 See V/9/ J/A+AS/7/1 See VIII/36/ J/A+AS/8/1 See VIII/11/ J/A+AS/9/1 See I/30/ J/A+AS/11/291 See VIII/36/ J/A+AS/14/1 See VIII/7A/ J/A+AS/14/557 See VIII/7A/ J/A+AS/16/71 See III/46/ J/A+AS/16/345 See I/2/ J/A+AS/16/445 See III/33/ J/A+AS/18/147 See VIII/36/ J/A+AS/18/463 See VI/30A/ J/A+AS/19/189 See II/26A/ J/A+AS/21/81 See II/63A/ J/A+AS/21/193 See III/34/ J/A+AS/22/239 See II/27/ J/A+AS/22/285 See V/13/ J/A+AS/23/283 See III/35/ J/A+AS/24/159 See II/29/ J/A+AS/25/213 See II/30/ J/A+AS/27/343 See II/20/ J/A+AS/27/443 See II/33/ J/A+AS/27/467 See I/63/ J/A+AS/28/123 See II/239/ J/A+AS/29/257 See III/40A/ J/A+AS/29/259 See II/35/ J/A+AS/30/1 See II/63A/ J/A+AS/30/261/ catalog.dat notes.dat (58Kb)

Ardeberg A., Maurice E.

Astron. Astrophys. Suppl. Ser. 30, 261 (1977)

Observations of supergiant stars in the Small Magellanic Cloud J/A+AS/30/349 See I/99/ J/A+AS/31/99 See VIII/43/ J/A+AS/31/243 See III/111/ J/A+AS/31/439/ table1.dat table3.dat table4.dat (58Kb)

Pellet A., Astier N., Viale A., Courtes G., Maucherat, A., Monnet, G., Simien, F. Astron. Astrophys. Suppl. Ser. 31, 439 (1978) A Survey of HII Regions in M31

J/A+AS/34/1 See II/51/ J/A+AS/34/477 See II/7A/ J/A+AS/35/23/ table1.dat (83Kb)

Altenhoff W.J., Downes D., Pauls T., Schraml J. Astron. Astrophys. Suppl. Ser. 35, 23 (1979) Survey of the Galactic Plane at 4.875 GHz

J/A+AS/36/95 See VIII/8A/ J/A+AS/36/163 See IV/1/ J/A+AS/36/237 See VIII/43/ J/A+AS/36/337 See V/18/ J/A+AS/37/575 See III/93/ J/A+AS/38/423 See V/35/ J/A+AS/38/449 See III/56/ J/A+AS/38/463 See II/14/ J/A+AS/39/39 See II/239/ J/A+AS/40/1 See II/57/ J/A+AS/40/47 See VIII/9/ J/A+AS/40/199 See V/14/ J/A+AS/40/307 See III/69A/ J/A+AS/41/151/ notes.dat table3.dat (78Kb)

Hummel E.

Astron. Astrophys. Suppl. Ser. 41, 151 (1980)

A radio continuum survey of bright galaxies at 1415 MHz. J/A+AS/41/339 See VIII/43/ J/A+AS/41/395 See II/58/ J/A+AS/41/421/ table2.dat table3.dat table4.dat table5.dat (30Kb)

Kotanyi C.G.

Astron. Astrophys. Suppl. Ser. 41, 421 (1980)

A radio continuum survey at 1.4 GHz of the galaxies in the Virgo region. J/A+AS/42/91 See II/76A/ J/A+AS/42/391/ table2 table3 (34Kb)

RULAND F., GRIFFIN R., GRIFFIN R., BIEHL D., HOLWEGER H.

Astron. Astrophys. Suppl. Ser. 42, 391 (1980)

Line Blocking and equivalent widths in the spectrum of Pollux J/A+AS/43/1/ notes.dat table3.dat (28Kb)

Fanti C., Mantovani F., Tomasi P.

Astron. Astrophys. Suppl. Ser., 43, 1-17 (1981)

New galactic sources in the BG catalogue: search at 1.4 GHz with the Westerbork synthesis radiotelescope. J/A+AS/43/27 See II/65/ J/A+AS/43/155 See VIII/43/ J/A+AS/43/259 See IV/9/ J/A+AS/43/265 See VI/14/ J/A+AS/44/137 See III/59/ J/A+AS/45/119 See II/66/ J/A+AS/45/207 See II/72/ J/A+AS/45/367 See VII/78/ J/A+AS/48/165 See II/78/ J/A+AS/48/485 See V/24/ J/A+AS/49/13 See I/99/ J/A+AS/49/99 See II/75/ J/A+AS/49/137 See VIII/47/ J/A+AS/49/427 See II/97/ J/A+AS/49/631 See VII/39A/ J/A+AS/49/715/ table2.dat table3.dat (54Kb)

Terzan A., Bijaoui A., Ju K.H., Ounnas C. Astron. Astrophys. Suppl. Ser. 49, 715 (1982) New variable stars in the direction of the bright cloud B in Sagittarius

J/A+AS/50/147 See I/74/ J/A+AS/52/63/ hicenter.fit (923Kb)

Burton W.B., Liszt H.S.

Astron. Astrophys. Suppl. Ser. 52, 63 (1983)

The gas distribution in the central region of the Galaxy. IV. A survey of neutral hydrogen in the region 349 l 13, -10 b 10, —v— 350 km/s

J/A+AS/52/237 See III/79/ J/A+AS/52/471 See III/102/ J/A+AS/52/479 See III/103/ J/A+AS/53/119 See II/134/ J/A+AS/53/219 See VII/54/ J/A+AS/53/223 See I/99/ J/A+AS/53/373 See VII/70A/ J/A+AS/54/55 See II/90/ J/A+AS/55/87 See I/99/ J/A+AS/55/371 See III/152/ J/A+AS/56/229 See II/100/ J/A+AS/56/363 See VII/71/ J/A+AS/57/213 See III/83/ J/A+AS/57/357 See II/182/ J/A+AS/57/385 See V/37/ J/A+AS/58/1/ table7.dat (84Kb)

Windhorst R.A., van Heerde G.M., Katgert P. Astron. Astrophys. Suppl. Ser. 58, 1 (1984) A deep Westerbork survey of areas with multicolor Mayall 4 m plates. I. The 1412 MHz catalogue, source counts and angular size statistics.

J/A+AS/58/39/  notes.dat table1.dat table4a.dat table4b.dat  (44Kb)

Windhorst R.A., Kron R.G., Koo D.C.

Astron. Astrophys. Suppl. Ser. 58, 39 (1984)

A deep Westerbork survey of areas with multicolor Mayall 4 m plates. II. Optical identifications. J/A+AS/58/197/ table3.dat (91Kb)

Reich W., Furst E., Steffen P., Reif K., Haslam C.G.T.

Astron. Astrophys. Suppl. Ser. 58, 197 (1984)

A radio continuum survey of the galactic plane at 11cm wavelength. I. The area 357.4 =l =76, -1.5 =b =1.5. J/A+AS/58/365 See VII/68A/ J/A+AS/58/565 See II/111/ J/A+AS/58/601 See V/49/ J/A+AS/58/735 See II/103/ J/A+AS/59/15 See III/105/ J/A+AS/59/145 See III/96/ J/A+AS/59/255 See VIII/37/ J/A+AS/59/461 See II/182/ J/A+AS/60/183 See II/127/ J/A+AS/60/339 See VI/29A/ J/A+AS/61/53 See III/104/ J/A+AS/61/331 See II/182/ J/A+AS/61/365 See II/133/ J/A+AS/61/451/ table3.dat (38Kb)

Walterbos R.A.M., Brinks E., Shane W.W. Astron. Astrophys. Suppl. Ser. 61, 451 (1985) A high resolution radio continuum survey of Messier 31. II. Observations at 1412 MHz, data reduction and the 37W catalogue of radio sources.

J/A+AS/61/473/ table1 (36Kb) Bernazzani M., Buonanno R., Corsi C.E., Hardy E. Astron. Astrophys. Suppl. Ser. 61, 473 (1985) Magnitudes, colors and positions for bright stars in the Large Magellanic Cloud bar

J/A+AS/62/355 See III/134/ J/A+AS/63/27 See VII/191/ J/A+AS/63/293 See IV/17A/ J/A+AS/64/21 See III/107/ J/A+AS/65/59 See III/108/ J/A+AS/65/195 See II/116/ J/A+AS/65/267/ table1.dat (48Kb)

Altschuler D.R.

Astron. Astrophys. Suppl. Ser. 65, 267 (1986)

A 5-GHz Survey of Radio Sources J/A+AS/65/349/ ref table1 table2 table3 table4 table5 table6 table7 (68Kb)

Petrosian A.R., Turatto M.

Astron. Astrophys. Suppl. Ser. 65, 349 (1986)

On the relation of Markarian galaxies with Zwicky clusters. I. Data J/A+AS/65/405 See II/182/ J/A+AS/65/485/ table1.dat (42Kb)

Roger R.S., Costain C.H., Stewart D.I.

Astron. Astrophys. Suppl. Ser. 65, 485 (1986) Spectral flux densities of radio sources at 22 MHz.

J/A+AS/65/561 See VIII/43/ J/A+AS/66/131 See I/121/ J/A+AS/66/181 See II/130/ J/A+AS/66/255 See VII/98A/ J/A+AS/67/327/ table1 (38Kb)

BUONANNO R., CORSI C.E., FERRARO I., FUSI PECCI F.

Astron. Astrophys. Suppl. Ser. 67, 327 (1987)

CCD photometry in globular clusters. II. NGC 7492. J/A+AS/68/113 See VI/37/ J/A+AS/68/117 See I/128/ J/A+AS/68/211 See II/182/ J/A+AS/69/87 See II/147/ J/A+AS/69/465 See III/124/ J/A+AS/70/77 See VIII/57/ J/A+AS/70/95/ table1.dat table2.dat table3.dat (61Kb)

de Grijp M.H.K., Miley G.K., Lub J.

Astron. Astrophys. Suppl. Ser. 70, 95 (1987)

Warm IRAS sources. I. A catalogue of AGN candidates from the point source catalog. J/A+AS/70/373 See III/154/ J/A+AS/70/481/ table1.dat (8Kb)

Bruch A., Fischer F.J., Wilmsen U.

Astron. Astrophys. Suppl. Ser. 70, 481 (1987) An atlas and catalogue of northern dwarf novae

J/A+AS/71/221/ table2.dat (54Kb) Oort M.J.A. Astron. Astrophys. Suppl. Ser. 71, 221 (1987) A deep WSRT 21 cm survey down to 0.1 mJy in the Lynx area.

J/A+AS/71/413 See II/122B/ J/A+AS/71/561 See II/182/ J/A+AS/72/259 See III/141/ J/A+AS/72/291 See V/71/ J/A+AS/73/37/ table1 (69Kb)

Linde P., Ardeberg A., Lindgren H., Lynga G. Astron. Astrophys. Suppl. Ser. 73, 37 (1988) B and V photometry of faint stars in the Large Magellanic Cloud

J/A+AS/73/103/ table2.dat (56Kb) Oort M.J.A., Steemers W.J.G., Windhorst R.A. Astron. Astrophys. Suppl. Ser. 73, 103 (1988) A deep 92cm survey of the Lynx area.

J/A+AS/73/195 See III/130/ J/A+AS/73/265 See VIII/43/ J/A+AS/74/83 See VII/159/ J/A+AS/74/315/ rad a262.dat (25Kb)

Righetti G., Giovannini G., Feretti L.

Astron. Astrophys. Suppl. Ser. 74, 315 (1988) Survey at 327 MHz of the A 262 cluster region.

J/A+AS/74/449 See I/146/ J/A+AS/75/67/ epol11cm.dat (40Kb)

Loiseau N., Reich W., Wielebinski R., Reich P., Munch W.

Astron. Astrophys. Suppl. Ser. 75, 67 (1988)

A survey of an area around the North Ecliptic Pole at λ11 cm wavelength. J/A+AS/75/93 See VI/47/ J/A+AS/75/441 See V/60/ J/A+AS/75/497 See VI/48/ J/A+AS/76/65 See II/144/

J/A+AS/76/101/  table3  (19Kb)

WESTERLUND B.E., GARNIER R., LUNDGREN K., PETTERSSON B., BREYSACHER J. Astron. Astrophys. Suppl. Ser. 76, 101 (1988) UBV and uvbyb photometry of stars in the region of the Zeta Sculptoris cluster

J/A+AS/76/205/  table3.dat table4.dat  (146Kb)

Terzan A., Ounnas CH. Astron. Astrophys. Suppl. Ser. 76, 205 (1988) A photometric study of the bright cloud B in Sagittarius. VI. 1592 new variable stars and 30 diffuse objects.

J/A+AS/77/31/  table2 table3 table4 table5 table6  (26Kb)

Vallee J.P., Roger R.S. Astron. Astrophys. Suppl. Ser. 77, 31 (1989) Radio surveys and source counts at 408 MHz and 1420 MHz towards the Abell 1314 cluster of galaxies

J/A+AS/77/447/  appendix table2  (62Kb)

REED B.C. Astron. Astrophys. Suppl. Ser. 77, 447 (1989) Photographic UBV photometry to V 21 in the Puppis window

J/A+AS/78/231/  table2.dat  (16Kb)

Wroblewski H., Torres C. Astron. Astrophys. Suppl. Ser. 78, 231 (1989) New proper-motion stars south of declination -40deg and right ascension between 00h00 and 04h30

J/A+AS/80/67 See III/161/

J/A+AS/80/103 See III/175/

J/A+AS/80/149/  table1.dat  (167Kb)

Wouterloot J.G.A., Brand J. Astron. Astrophys. Suppl. Ser. 80, 149 (1989) IRAS sources beyond the solar circle. I. CO observations

J/A+AS/80/215/  radlunar.dat  (41Kb)

Akujor C.E., Joshi M.N., Kazes I. Astron. Astrophys. Suppl. Ser. 80, 215 (1989) Meter wavelength structures, flux densities and accurate positions of weak radio sources.

J/A+AS/80/285 See VI/34/

J/A+AS/80/299 See VII/119/

J/A+AS/81/221 See III/89/

J/A+AS/81/401 See II/164/

J/A+AS/81/421 See II/159/

J/A+AS/82/41/  table1.dat table3.dat  (58Kb)

Kulkarni V.K., Mantovani F., Pauliny-Toth I.I.K. Astron. Astrophys. Suppl. Ser. 82, 41 (1990) Spectral index distribution for a sample of sources from the B3 catalogue.

J/A+AS/82/71/  table3 table4  (132Kb)

Ortolani S.,Gratton R. Astron. Astrophys. Suppl. Ser. 82, 71 (1990) Deep photometry of globular clusters. XII. The metal-poor clusters NGC 5694 and Terzan 8

J/A+AS/82/145 See II/189/

J/A+AS/82/279 See VIII/43/

J/A+AS/82/331 See II/163/

J/A+AS/82/391 See VII/136/

J/A+AS/83/317/  table1  (16Kb)

Wroblewski H., Torres C. Astron. Astrophys. Suppl. Ser. 83, 317 (1990) New proper motion determination of Luyten catalogue stars (LTT) south of declination -40 and right ascension between 00h 30m.

J/A+AS/83/539/  table2  (94Kb)

Reich W., Reich P., Furst E Astron. Astrophys. Suppl. Ser. 83, 539 (1990) The Effelsberg 21cm radio continuum survey of the Galactic plane between l=357deg and l=95.5deg

J/A+AS/84/59/  table2a table2b  (453Kb)

FERRARO F.R., CLEMENTINI G., FUSI PECCI F., BUONANNO R., ALCAINO G. Astron. Astrophys. Suppl. Ser. 84, 59 (1990) CCD-photometry of the galactic globular cluster NGC 2808.

J/A+AS/84/133 See II/166/

J/A+AS/84/527 See VII/183/

J/A+AS/85/805/  table1.dat  (647Kb)

Fuerst E., Reich W., Reich P., Reif K. Astron. Astrophys. Suppl. Ser. 85, 805 (1990) A radio continuum survey of the galactic plane at 11cm wavelength. IV. A list of 6483 small diameter sources.

J/A+AS/85/839 See II/165/

J/A+AS/85/885 See III/160A/

J/A+AS/85/915/  table  (314Kb)

DE GEUS E.J., LUB J., VAN DE GRIFT E. Astron. Astrophys. Suppl. Ser. 85, 915 (1990) Walraven photometry of nearby southern OB associations

J/A+AS/85/989 See II/167/

J/A+AS/85/1069 See V/69/

J/A+AS/85/1179 See V/59/

J/A+AS/86/107 See II/158A/

J/A+AS/87/1/  table1.dat  (90Kb)

Parijskij Yu.N., Bursov N.N., Lipovka N.M., Soboleva N.S., Temirova A.V. Astron. Astrophys. Suppl. Ser. 87, 1 (1991) The RATAN-600 7.6 cm catalog of radio sources from ”Experiment Cold-80”.

J/A+AS/87/159 See I/159A/

J/A+AS/87/481 See II/170/

J/A+AS/88/277 See III/155/

J/A+AS/88/355/  table1 table2  (13Kb)

GREEN D.A. Astron. Astrophys. Suppl. Ser. 88, 355 (1991) Compact radio sources in the galactic plane near l=140deg, b=0deg

J/A+AS/89/311/  rad strs.dat  (29Kb)

Requieme Y., Mazurier J.M. Astron. Astrophys. Suppl. Ser. 89, 311 (1991) Optical positions of 221 radio stars obtained with the Bordeaux automatic meridian circle.

J/A+AS/89/415 See II/182/

J/A+AS/89/429 See III/162A/

J/A+AS/89/469 See III/148A/

J/A+AS/90/327/  catalog.dat  (287Kb)

te Lintel Hekkert P., Caswell J.L., Habing H.J., Haynes R.F., Norris R.P. Astron. Astrophys. Suppl. Ser. 90, 327 (1991) 1612 MHz OH survey of IRAS point sources I. Observations made at Dwingeloo, Effelsberg and Parkes

J/A+AS/90/451/  table2  (96Kb)

Terzan A., Gosset E. Astron. Astrophys. Suppl. Ser. 90, 451-493 (1991) A photometric study of the bright cloud B in Sagittarius. VII. 1165 new variable stars and 65 diffuse objects.

J/A+AS/91/129/  table2.dat  (34Kb)

Wroblewski H., Torres C. Astron. Astrophys. Suppl. Ser. 91, 129 (1991) New proper-motion stars south of declination -40deg and right ascension between 04h30 and 16h00

J/A+AS/91/171/  table4  (83Kb)

DEMERS S., IRWIN M.J. Astron. Astrophys. Suppl. Ser. 91, 171 (1991) Catalogue of blue stars between the Magellanic Clouds

J/A+AS/91/285/  table2.dat table3.dat table4.dat table5.dat  (107Kb)

Terlevich R., Melnick J., Masegosa J., Moles M., Copetti M.V.F. Astron. Astrophys. Suppl. Ser. 91, 285 (1991) A spectrophotometric catalogue of HII galaxies.

J/A+AS/91/337/  table2  (13Kb)

Soboleva N.S., Temirova A.V., Reich. P. Astron. Astrophys. Suppl. Ser. 91, 337 (1991) A list of steep spectrum radio spectra sources

J/A+AS/91/371/  table1 table2 table3  (107Kb)

Paturel G., Petit C., Kogoshvili N., Dubois P., Bottinelli L., Fouque P., Garnier R., Gouguenheim L. Astron. Astrophys. Suppl. Ser. 91, 371 (1991) An Extragalactic Database. IV. Errors and misprints in catalogues of galaxies.

J/A+AS/92/1/  stars.dat table1.dat table3.dat table5.dat  (270Kb)

Glushneva I.N., Kharitonov A.V., Knyazeva L.N., Shenavrin V.I. Astron. Astrophys. Suppl. Ser. 92, 1 (1992) Secondary Spectrophotometric standards

J/A+AS/92/43/  table2.dat  (28Kb)

Lindqvist M., Winnberg A., Habing H.J., Matthews H.E. Astron. Astrophys. Suppl. Ser. 92, 43 (1992) OH/IR stars close to the Galactic centre. I. Observational data

J/A+AS/92/449/  table1  (24Kb)

Wroblewski H., Torres C. Astron. Astrophys. Suppl. Ser. 92, 449 (1992) New proper motion determination of Luyten catalogue stars (LTT) south of declination -40 and right ascension between 04h30 and 16h00m

J/A+AS/92/625/  table2.dat  (208Kb)

Walterbos R.A.M., Braun R. Astron. Astrophys. Suppl. Ser. 92, 625 (1992) The interstellar medium of M 31. III. Narrow-band imagery in H-alpha and SII.

J/A+AS/92/685/  table2a table2b table2c  (20Kb)

Altamore A., Rossi C., Viotti R., Baratta G.B. Astron. Astrophys. Suppl. Ser. 92, 685-719 (1992) The spectrum of the VV Cep star KQ Puppis (Boss 1985). II. Atlas of the optical and ultraviolet spectrum

J/A+AS/92/841/  table1  (41Kb)

Knude J. Astron. Astrophys. Suppl. Ser. 92, 841-861 (1992) uvby-beta observations of 528 type B stars with V between the 8th and 9th magnitude.

J/A+AS/93/61/  table1 table2 table4  (50Kb)

Ruelas-Mayorga R.A., Teague P.F. Astron. Astrophys. Suppl. Ser. 93, 61 (1992) Distribution and studies of the infrared stellar population in the galaxy. III. Baade’s window photometry.

J/A+AS/93/121/  table4 table6  (57Kb)

NYMAN L.A., BOOTH R.S., CARLSTROEM U., HABING H.J., HESKE A., SAHAI R., STARK R., VAN DER VEEN W.E.C.J., WINNBERG A. Astron. Astrophys. Suppl. Ser. 93, 121 (1992) A survey of circumstellar CO emission from a sample of IRAS sources

J/A+AS/93/151 See II/177/

J/A+AS/93/189/  table2 table3 table4 table5 table6 table7  (41Kb)

RODRIGUEZ E., ROLLAND A., LOPEZ DE COCA P., GARCIA-LOBO E., SEDANO J.L. Astron. Astrophys. Suppl. Ser. 93, 189 (1992) Stromgren photometry of three double mode high amplitude Delta Scuti stars

J/A+AS/93/255/  table1 table2 table4  (34Kb)

GUTHRIE B.N.G. Astron. Astrophys. Suppl. Ser. 93, 255 (1992) Axial ratios of edge-on spirals

J/A+AS/93/293/  table6  (302Kb)

TUCHOLKE H.J. Astron. Astrophys. Suppl. Ser. 93, 293 (1992) Astrometry of globular clusters. I. Relative proper motions for stars in 47 Tuc.

J/A+AS/93/311/  table6  (51Kb)

TUCHOLKE H.J. Astron. Astrophys. Suppl. Ser. 93, 311 (1992) Astrometry of globular clusters. II. Relative proper motions and the space motion of NGC 362

J/A+AS/93/373/  table1.dat  (25Kb)

Franco G.A.P Astron. Astrophys. Suppl. Ser. 93, 373 (1992) Four-colour uvby and H-beta photometry of all stars earlier than G0 and brighter than 10.7mag in Selected Area 203 towards the Chamaeleon-Musca dark clouds

J/A+AS/93/383/  tables  (19Kb)

MUNARI U., YUDIN B.F., TARANOVA O.G., MASSONE G., MARANG F., ROBERTS G., WINKLER H., WHITELOCK P.A. Astron. Astrophys. Suppl. Ser. 93, 383 (1992) UBVRIJHKL photometric catalogue of symbiotic stars

J/A+AS/93/399/  table1 table2  (23Kb)

WIERINGA M.H., KATGERT P. Astron. Astrophys. Suppl. Ser. 93, 399 (1992) Optical identifications of faint ultra-steep spectrum radio sources

J/A+AS/93/463/  table1.dat  (16Kb)

Bruch A., Meijer J., Naumann M., Schimpke T., Ungruhe R., Vogt N. Astron. Astrophys. Suppl. Ser. 93, 463 (1992) An astrometric catalogue of southern and equatorial dwarf novae

J/A+AS/93/553/  table1  (13Kb)

SAURER W., PFITSCHER K., WEINBERGER R., HARTL H. Astron. Astrophys. Suppl. Ser. 93, 553 (1992) Photoelectric UBV photometry of stars in four fields near the galactic plane.

J/A+AS/93/557/  tables  (6Kb)

PRIETO M., LONGLEY D.P.T., PEREZ E., BECKMAN J.E., VARELA A.M., CEPA J. Astron. Astrophys. Suppl. Ser. 93, 557 (1992) Multiband analysis of the surface brightness distributions of Sb and Sc spirals CCD images. I. Observations.

J/A+AS/94/73/  ngc2421 ngc2439 ngc2489 ngc2567 ngc2627 ngc2658 ngc2910  (57Kb)

RAMSAY G., POLLACO D.L. Astron. Astrophys. Suppl. Ser. 94, 73 (1992) (Erratum note in Astron. Astrophys. Suppl. Ser. 104, 377, 1994) CCD observations in 7 open clusters: NGC 2421,2439,2489,2567,2627,2658,2910

J/A+AS/94/327/  tables.dat  (76Kb)

Trevese D., Flin P., Migliori L., Hickison P., Pittella G. Astron. Astrophys. Suppl. Ser. 94, 327 (1992) Properties of nearby clusters of galaxies. I. A 195, A 465, A 1185, A 1213, A 1413, A 1775, A 2319 and A 2597

J/A+AS/94/399/  refs.dat table1.dat table2.dat table4.dat  (224Kb)

Cahn J.H., Kaler J.B., Stanghellini L. Astron. Astrophys. Suppl. Ser. 94, 399 (1992) A catalogue of absolute fluxes and distances of planetary nebulae.

J/A+AS/94/519/  table1  (6Kb)

JORDI C., TRULLOLS E., ROSSELLO G., LAHULLA F. Astron. Astrophys. Suppl. Ser. 94, 519-525 (1992) A photometric analysis of the Cepheus OB3 region

J/A+AS/95/1/  table1 table2 table3 table4  (13Kb)

Gregorini L., Vettolani G., de Ruiter H.R., Parma P. Astron. Astrophys. Suppl. Ser. 95, 1 (1992) Sample of dumb-bell galaxies in southern rich clusters

J/A+AS/95/273 See III/165/

J/A+AS/95/355/  catalog  (100Kb)

CARRASCO G., LOYOLA P. Astron. Astrophys. Suppl. Ser. 95, 355 (1992) Santiago Fundamental Catalogue. A catalogue of 1105 FK5 stars (equinox J2000.0)

J/A+AS/96/269/  ( total of 7 files)  (627Kb)

SCHALLER G., SCHAERER D., MEYNET G., MAEDER A.

Astron. Astrophys. Suppl. Ser. 96, 269 (1992) New grids of stellar models from 0.8 to 120 Msolar at Z=0.020 and Z=0.001

J/A+AS/96/379 See II/208/

J/A+AS/96/593 See III/196/

J/A+AS/97/435/  table1  (28Kb)

Quiniento Z.M., Cersosimo J.C. Astron. Astrophys. Suppl. Ser., 97, 435-442 (1993) Radio spectra of quasars. III.

J/A+AS/97/443/  table2 table3  (443Kb)

Kaastra J.S., Mewe R. Astron. Astrophys. Suppl. Ser. 97, 443 (1993) X-ray emission from thin plasmas. I. Multiple Auger ionisation and fluorescence processes for Be to Zn.

J/A+AS/97/483 See III/175/

J/A+AS/97/587/  table1 table2  (15Kb)

Ruelas-Mayorga R.A., Teague P.F. Astron. Astrophys. Suppl. Ser. 97, 587 (1993) Distribution and studies of the infrared stellar population in the Galaxy. V. Other clear regions around the galactic centre.

J/A+AS/97/603/  table4.dat  (104Kb)

Rebeirot E., Azzopardi M., Westerlund B.E. Astron. Astrophys. Suppl. Ser. 97, 603 (1993) Carbon Stars in the Small Magellanic Cloud. II. Catalogue of 1707 objects with identifications and spectrophotometry.

J/A+AS/97/729/  appendix  (49Kb)

Le Bertre T. Astron. Astrophys. Suppl. Ser. 97, 729 (1993) Oxygen-rich late-type star lightcurves in the 1-20microm. range.

J/A+AS/97/807 See VI/57/

J/A+AS/97/835/  tables  (83Kb)

FLYNN C., FREEMAN K.C. Astron. Astrophys. Suppl. Ser. 97, 835 (1993) A catalogue of K giants at the South Galactic Pole: broadband and DDO photometry and radial velocities

J/A+AS/97/951/  table4  (46Kb)

Schuster W.J., Parrao L., Contreras Martinez M.E. Astron. Astrophys. Suppl. Ser. 97, 951 (1993) uvby-beta photometry of high-velocity and metal-poor stars. VI. A second catalogue, and stellar populations of the Galaxy

J/A+AS/98/29 See VII/184/

J/A+AS/98/393/ refs.dat table1.dat table2.dat table3.dat table4.dat table5.dat table6.dat table7.dat (53Kb) Stickel M., Fried J.W., Kuehr H. Astron. Astrophys. Suppl. Ser. 98, 393 (1993) The complete sample of 1 Jy BL Lac Objects. II. Observational Data.

J/A+AS/98/505/  table1 table2  (30Kb)

Hoekzemal N.M., Lamers H.J.G.L.M., van Genderen A.M. Astron. Astrophys. Suppl. Ser. 98, 505 (1993) Walraven photometry of stars near the Luminous Blue Variable AG Carinae

J/A+AS/98/523/ table1 table2 table3 table4 table5 table6 table7 table8 table9 table10 table11 table12 table13 table14 table15 table16 table17 table18 table19 table20 table21 table22 (304Kb) SCHAERER D., MEYNET G., MAEDER A. SCHALLER G. Astron. Astrophys. Suppl. Ser. 98, 523 (1993) Grids of stellar models. II. From 0.8 to 120 Msolar at Z=0.008

J/A+AS/98/589/  table1 table2 table3 table4 table5 table6 table7 table8  (385Kb)

Wouterloot J.G.A., Brand J., Fiegle K. Astron. Astrophys. Suppl. Ser. 98, 589 (1993) IRAS Sources behind the Solar circle. III. Observations of H2O, OH, CH3OH and CO

J/A+AS/99/31/ table3a table3b table4a table4b table5a table5b table6a table6b table7a table7b (201Kb)

GUGLIELMO F., EPCHTEIN N., LE BERTRE T., FOUQUE P., HRON J., KERSCHBAUM F., LEPINE J.R.D. Astron. Astrophys. Suppl. Ser. 99, 31 (1993) Identification of 106 new infrared carbon stars in the IRAS Point Source Catalog: Near-infrared photometry and their space distribution in the Galaxy

J/A+AS/99/291/ refs.dat table1.dat table2.dat table3.dat table5a.dat table5b.dat table5c.dat tables.tex (682Kb) Loup C., Forveille T., Omont A., Paul J.F. Astron. Astrophys. Suppl. 99, 291 (1993) CO and HCN observations of circumstellar envelopes. A catalogue - mass loss rates and distributions

J/A+AS/99/545/  miyun232.dat  (46Kb)

Zhang X., Zhen Y., Chen H., Wang S. Astron. Astrophys. Suppl. Ser. 99, 545 (1993) The Miyun 232 MHz survey. I. Fields centred at RA: 00h41m, Dec: 41^{∘}12’ and RA: 07h00m, Dec: 35^{∘}00’

J/A+AS/99/591/  table7  (76Kb)

Gochermann J., Grothues H.-G., Oestreicher M.O., Berghoefer T., Schmidt-Kaler T. Astron. Astrophys. Suppl. Ser. 99, 591 (1993) UBV photometry of galactic foreground and LMC member stars. I. Galactic foreground stars

J/A+AS/100/47/  table1.dat table2.dat  (243Kb)

Garcia A.M. Astron. Astrophys. Suppl. Ser. 100, 47 (1993) General study of group membership. II. Determination of nearby groups.

J/A+AS/100/173 See V/76/

J/A+AS/100/243/  table6.dat  (75Kb)

Zhao J.L., Tian K.P., Pan R.S., He Y.P., Shi H.M. Astron. Astrophys. Suppl. Ser. 100, 243 (1993) Study of proper motions in the region of the open cluster M67 and membership of stars

J/A+AS/100/395 See III/175/

J/A+AS/100/473/  table6 table7  (26Kb)

ASSENDORP R., WESSELIUS P.R. Astron. Astrophys. Suppl. Ser. 100, 473 (1993) IRAS pointed observations data processing

J/A+AS/100/489/  table1.dat table2.dat  (17Kb)

Carrasco G., Loyola P. Astron. Astrophys. Suppl. Ser. 100, 489 (1993) UBVRI photometry of FKSZ stars. IV.

J/A+AS/100/517/  table1  (11Kb)

VAN DESSEL A., SINACHOPOULOS D. Astron. Astrophys. Suppl. Ser. 100, 517 (1993) CCD astrometry and instrumental Delta-V photometry of wide visual double stars. III. Differential measurements of often observed southern pairs.

J/A+AS/100/571/  table3 table4 table5 table6 table7 table8 table9 table10 table11 table12  (58Kb)

Rodriguez E., Rolland A., Lopez de Coca P. Astron. Astrophys. Suppl. Ser. 100, 571 (1993) Simultaneous uvby-beta photometry of SX Phe stars

J/A+AS/100/583/  table2 table3 table4 table5 table6 table7 table8  (40Kb)

Boisson C., Durret F., Balkowski C., Proust D. Astron. Astrophys. Suppl. Ser. 100, 583 (1993) Infrared and optical photometry of galaxies in four clusters and of a sample of early-type galaxies.

J/A+AS/100/591 See II/182/

J/A+AS/100/607/  table4  (29Kb)

HAMBLY N.C., HAWKINS M.R.S., JAMESON R.F. Astron. Astrophys. Suppl. Ser. 100, 607 (1993) Very low mass proper motion members in the Pleiades

J/A+AS/100/647 See VI/96/

J/A+AS/101/1/  table1 table2 table3 table4  (267Kb)

EKBERG J. O.

Astron. Astrophys. Suppl. Ser. 101, 1 (1993) Wavelengths and transition probabilities of the 3d6-3d54p and 3d54s-3d54p transition arrays of Fe III

J/A+AS/101/49/  table3  (15Kb)

LIU Q. Astron. Astrophys. Suppl. Ser. 101, 49 (1993) Four-colour Photometric Study of the short-period eclipsing binary V Crt

J/A+AS/101/59/  notes.dat refs.dat table1.dat table23.dat  (161Kb)

Molinari S., Liseau R., Lorenzetti D. Astron. Astrophys. Suppl. Ser. 101, 59 (1993) The Exciting Sources of Herbig-Haro objects. I. A Catalogue of 1-20um observations

J/A+AS/101/87/ bv-cen.dat hl-aqr.dat v426oph.dat v603aql.dat v2051oph.dat v3885sgr.dat vy-scl.dat ww-cet.dat (1.0 Mb) Hollander A., Kraakman H., van Paradijs J. Astron. Astrophys. Suppl. Ser., 101, 87 (1993) Walraven photometry of eight cataclysmic variables

J/A+AS/101/153/  table1 table5  (105Kb)

Palagi F., Cesaroni R., Comoretto G., Felli M., Natale V. Astron. Astrophys. Suppl. Ser. 101, 153 (1993) Classification and statistical properties of galactic H2O masers

J/A+AS/101/195/  table4-7 table8  (54Kb)

PIERSIMONI A.M., DI PAOLANTONIO A., BURCHI R., DE SANTIS R. Astron. Astrophys. Suppl. Ser. 101, 195 (1993) Photoelectric Photometry of field variables. II

J/A+AS/101/259/  table table.tex  (16Kb)

Galli M., Cappi A., Focardi P., Gregorini L., Vettolani G. Astron. Astrophys. Suppl. Ser. 101, 259 (1993) Redshifts of southern rich clusters

J/A+AS/101/415/ table1 table2 table3 table4 table5 table6 table7 table8 table9 table10 table11 table12 table13 table14 table15 table16 table17 table18 table19 table20 table21 table22 (304Kb) CHARBONNEL C., MEYNET G., MAEDER A., SCHALLER G., SCHAERER D. Astron. Astrophys. Suppl. Ser. 101, 415 (1993) Grids of stellar models. III. From 0.8 to 120 Msolar at Z=0.004

J/A+AS/101/421/  table1 table2  (16Kb)

RODRIGUEZ E., ROLLAND A., LOPEZ DE COCA P. Astron. Astrophys. Suppl. Ser. 101, 421 (1993) Simultaneous uvby-beta photometry of GP Andromedae

J/A+AS/101/475/  table1  (8Kb)

Quintana H., de Souza R. Astron. Astrophys. Suppl. Ser. 101, 475 (1993) Spectroscopic observations of the galaxy cluster A 3571 (SC 1344-325)

J/A+AS/101/485/  table1  (46Kb)

BOUVIER J., CABRIT S., FERNANDEZ M., MARTIN E.L., MATTHEWS J.M. Astron. Astrophys. Suppl. Ser. 101, 485 (1993) COYOTES I. Multisite UBVRI photometry of 24 pre-main-sequence stars of the Taurus-Auriga cloud

J/A+AS/101/513/  table1.dat table2.dat  (12Kb)

Groenewegen M.A.T., de Jong T., Baas F. Astron. Astrophys. Suppl. Ser. 101, 513 (1993) Near-infrared and sub-millimeter photometry of carbon stars

J/A+AS/101/521 See III/175/

J/A+AS/101/551/  table1.dat  (65Kb)

Leroy J.L. Astron. Astrophys. Suppl. Ser. 101, 551 (1993) Optical polarization of 1000 stars within 50 pc from the Sun

J/A+AS/101/563/  table1  (125Kb)

CLAUSEN J.V., GARCIA J.M., GIMENEZ A., HELT B.E., VAZ L.P.R. Astron. Astrophys. Suppl. Ser. 101, 563 (1993) Four-colour photometry of eclipsing binaries. XXXV. Light curves of GG Lupi: Young metal-deficient B stars.

J/A+AS/102/11/  table2 table3  (21Kb)

NOEL F. Astron. Astrophys. Suppl. Ser. 102, 11 (1993) Observations of the Sun during 1990-1992 with the astrolabe of Santiago

J/A+AS/102/57/  notes table5  (33Kb)

Bottinelli L., Durand N., Fouque P., Garnier R., Gouguenheim L., Loulergue M., Paturel G., Petit C., Teerikorpi P. Astron. Astrophys. Suppl. Ser. 102, 57 (1993) Observational data for the kinematics of the local universe. II. Second set of radial velocity measurements.

J/A+AS/102/79 See II/188/

J/A+AS/102/89/  table8 table10 table12 table18  (1.2 Mb)

OLSEN E.H. Astron. Astrophys. Suppl. Ser. 102, 89 (1993) Stroemgren four-colour uvby photometry of G5-type HD stars brighter than mv=8.6

J/A+AS/102/251/  table3.dat table4.dat  (67Kb)

Meyssonnier N., Lequeux J., Azzopardi M. Astron. Astrophys. Suppl. Ser. 102, 251 (1993) An objective-prism survey of emission-line objects in M 31.

J/A+AS/102/269/  table1.dat  (201Kb)

NAVE G., JOHANSSON S. Astron. Astrophys. Suppl. Ser. 102, 269 (1993) Highly-excited levels of Fe I obtained from laboratory and solar Fourier transform and grating spectra. II. Laboratory and solar identifications

J/A+AS/102/339/ table1 table2 table3 table4 table5 table6 table7 table8 table9 table10 table11 table12 table13 table14 table15 table16 table17 table18 table19 table20 table21 table22 (284Kb) SCHAERER D., CHARBONNEL C., MEYNET G., MAEDER A., SCHALLER G. Astron. Astrophys. Suppl. Ser. 102, 339 (1993) Grids of stellar models. IV. From 0.8 to 120 M_{⊙} at Z=0.040

J/A+AS/102/343/  table3 table4 table5a table6  (39Kb)

SCALTRITI F., PIIROLA V., COYNE G.V., KOCH S.J.R.H., ELIAS N.M., HOLENSTEIN B.D. Astron. Astrophys. Suppl. Ser. 102, 343 (1993) UBVRI linear and circular polarization of RS CVn-type binaries

J/A+AS/102/361/  table1.dat  (14Kb)

Boffin H.M.J., Abia C., Isern J., Rebolo R. Astron. Astrophys. Suppl. Ser. 102, 361 (1993) A Catalogue of Li abundances and equivalent widths in a sample of galactic C-stars

J/A+AS/102/397/  table1.dat  (114Kb)

Guarnieri M. D., Bragaglia A., Fusi Pecci F. Astron. Astrophys. Suppl. Ser. 102, 397 (1993) Colour magnitude diagram for the globular cluster M13

J/A+AS/102/401 See III/174/

J/A+AS/102/435/  table1 table2 table3 table4 table5 table6 table7 table8 table9  (38Kb)

BIEMONT E., QUINET P., ZEIPPEN C.J. Astron. Astrophys. Suppl. Ser. 102, 435 (1993) Delta(n) = 2 allowed transitions of neutral sulphur within the visible and infrared spectral ranges

J/A+AS/102/451/  notes table2  (116Kb)

Meyssonnier N., Azzopardi M. Astron. Astrophys. Suppl. Ser. 102, 451 (1993) A new catalogue of H-alpha emission-line stars and small nebulae in the Small Magellanic Cloud

J/A+AS/102/603/  table1a table1b table2 table3  (141Kb)

Edvardsson B., Andersen J., Gustafsson B., Lambert D.L., Nissen P.E., Tomkin J. Astron. Astrophys. Suppl. Ser. 102, 603 (1993) The chemical evolution of the galactic disk II. Observational data.

J/A+AS/102/607/  table1.dat  (6Kb)

Puric J., Djenize S., Sreckovic A., Bukvic S., Pivalica S., Labat J. Astron. Astrophys. Suppl. Ser. 102, 607 (1993) Stark widths of singly-ionized iron spectral lines

J/A+AS/102/611/  table2 title  (415Kb)

Steppe H., Paubert G., Sievers A., Reuter H.P., Greve A., Liechti S., Le Floch B., Brunswig W., Menedez C., Sanchez S. Astron. Astrophys. Suppl. Ser. 102, 611 (1993) Millimeter continuum measurements of extragalactic radio sources. III.

J/A+AS/102/643/  table1  (42Kb)

MULLER P. Astron. Astrophys. Suppl. Ser. 102, 643 (1993) Mesures d’etoiles doubles faite a Nice (4e serie)

J/A+AS/103/1/  table1.dat table2.dat table3.dat table4.dat  (27Kb)

Adelman S.J., Knox Jr J.R. Astron. Astrophys. Suppl. Ser. 103, 1 (1994) ubvy photometry of the magnetic chemically peculiar stars 63 Andromedae, HD 192913, HR 8240, and 108 Aquarii

J/A+AS/103/57/  table1a table1b table2a table2b table3a table3b  (23Kb)

MIHAJLOV A.A., DIMITRIJEVIC M.S., IGNJATOVIC LJ., DJURIC Z. Astron. Astrophys. Suppl. Ser. 103, 57 (1994) Spectral coefficients of emission and absorption due to ion-atom radiation collisions in solar atmosphere

J/A+AS/103/67/  lmc.dat smc.dat  (67Kb)

de Loore C., Vanbeveren D. Astron. Astrophys. Suppl. Ser. 103, 67 (1994) Evolutionary sequences for massive close binary stars with Magellanic Cloud metallicities with Rogers-Iglesias opacities

J/A+AS/103/83/  spec-a.dat spec-b.dat spec-c.dat table1.dat  (225Kb)

Borovicka J. Astron. Astrophys. Suppl. Ser. 103, 83 (1994) Line identifications in a fireball spectrum

J/A+AS/103/97/ table1 table2 table3 table4 table5 table6 table7 table8 table9 table10 table11 table12 table13 table14 table15 table16 table17 table18 table19 table20 table21 table22 table23 table24 table25 table26 table27 table28 table29 table30 table31 table32 table33 (453Kb) MEYNET G., MAEDER A., SCHALLER G., SCHAERER D., CHARBONNEL C. Astron. Astrophys. Suppl. Ser. 103, 97 (1994) Grids of massive stars with high mass loss rates. V. From 12 to 120 Mo at Z=0.001, 0.004, 0.008, 0.020 and 0.040

J/A+AS/103/183/  table3 table3.tex  (203Kb)

JETSU L., TUOMINEN I., ANTOV A., BOPP B.W., CHUGAINOV P.F., GRANKIN K.N., LINNALUOTO S., MELNIKOV S.YU., NATIONS H.L., PANOV K.P., SHEVCHENKO V.S. Astron. Astrophys. Suppl. Ser. 103, 183 (1994) Collected photometry of FK Comae Berenices

J/A+AS/103/201 See III/174/

J/A+AS/103/245/  table3 table4 table5  (60Kb)

MOLINARI E., BANZI M., BUZZONI A., CHINCARINI G., PEDRANA M.D. Astron. Astrophys. Suppl. Ser. 103, 245 (1994) Multicolor Photometry of Clusters of Galaxies: A3284, A3305 and A1942

J/A+AS/103/273/ table2.dat table3.dat table4.dat table5.dat table6.dat table7.dat table8.dat table9.dat table10.dat table11.dat (19Kb) Lennon D.J., Burke V.M. Astron. Astrophys. Suppl. Ser. 103, 273 (1994) Atomic Data from the IRON Project. II. Effective collision strengths for infrared transitions in carbon-like ions

J/A+AS/103/293/  table1 table2  (37Kb)

VELA VILLAHOZ E., SANCHEZ ALMEIDA J., WITTMANN A.D. Astron. Astrophys. Suppl. Ser. 103, 293 (1994) Spectral lines unaffected by instrumental polarization. II. Selected lines of astrophysical interest.

J/A+AS/103/349 See III/175/

J/A+AS/103/427/  table2,dat table3,dat  (45Kb)

KOLESNIK Y.B., DAVILA H. Astron. Astrophys. Suppl. Ser. 103, 427 (1994) The Second Quito Astrolabe Catalogue

J/A+AS/103/445/  table2  (33Kb)

LEUENHAGEN U., HEBER U., JEFFERY C.S. Astron. Astrophys. Suppl. Ser. 103, 445 (1994) High-resolution optical spectroscopy of the hot R CrB star V348 Sgr

J/A+AS/103/541/  refs.dat table1.dat  (0Kb)

Brand J., Cesaroni R., Caselli P., Catarzi M., Codella C., Comoretto G., Curioni G.P., Curioni P., di Franco S., Felli M., Giovanardi C., Olmi L., Palagi F., Palla F., Panella D., Pareschi G., Rossi E., Speroni N., Tofani G. Astron. Astrophys. Suppl. Ser. 103, 541 (1994) The Arcetri Catalogue of H2O Maser Sources Update

J/A+AS/103/573/  table1a table1b table2a table2b table3 table4  (158Kb)

POULAIN P., NIETO J.-L. Astron. Astrophys. Suppl. Ser. 103, 573 (1994) UBVRI photoelectric photometry of bright southern early-type galaxies

J/A+AS/104/1/  table1  (18Kb)

FAUNDEZ-ABANS M., CUEVAS H., HERTLING G. Astron. Astrophys. Suppl. Ser. 104, 1 (1994) Search for faint ring-shaped galaxies in the -77 deg to -87 deg declination interval

J/A+AS/104/9/  table1  (64Kb)

FRANCO G.A.P. Astron. Astrophys. Suppl. Ser. 104, 9 (1994) uvby-beta photometry of all stars earlier than G0 and brighter than m_{pg} = 10.3 mag in Selected Area 194

J/A+AS/104/89/  table2 table3 table4 table5 table6 table7  (92Kb)

POTEKHIN A.Y., VARSHALOVICH D.A. Astron. Astrophys. Suppl. Ser. 104, 89 (1994) Non-variability of the fine-structure constants over cosmological time scale

J/A+AS/104/161/  table3  (60Kb)

TUCHOLKE H.-J., SCHOLZ R.-D., BROSCHE P. Astron. Astrophys. Suppl. Ser. 104, 161 (1994) Proper motion study of the globular cluster M 3

J/A+AS/104/169/  table1  (21Kb)

PHILLIPS J.P., CUESTA L. Astron. Astrophys. Suppl. Ser. 104, 169 (1994) A near infrared survey of northern planetary nebulae

J/A+AS/104/179/  ( total of 7 files)  (3.0 Mb)

GOUDFROOIJ P., HANSEN L., JOERGENSEN H.E., NOERGAARD-NIELSEN H.U., DE JONG T., VAN DEN HOEK L.B. Astron. Astrophys. Suppl. Ser. 104, 179 (1994) Interstellar matter in Shapley-Ames elliptical galaxies. I. Multicolour CCD surface photometry

J/A+AS/104/233/  table2  (21Kb)

PETTERSSON B., REIPURTH B. Astron. Astrophys. Suppl. Ser. 104, 233 (1994) Young stars associated with the Vela Molecular Ridge. I. VMR clouds C and D, Collinder 197 and Vela R2.

J/A+AS/104/271/  table2 table3  (19Kb)

Gavazzi G., Garilli B., Carrasco L., Boselli A., Cruz-Gonzalez I. Astron. Astrophys. Suppl. Ser. 104, 271 (1994) A CCD survey of galaxies. II: observations with the 2.1 m telescope at San Pedro Martir

J/A+AS/104/315/ liste2.txt liste3.dat table1.dat table2.dat table3.dat table4a.dat table4b.dat table5.dat table6.dat (101Kb) The P.S., de Winter D., Perez M.R. Astron. Astrophys. Suppl. Ser. 104, 315 (1994) A new catalogue of members and candidate members of the Herbig Ae/Be (HAEBE) stellar group

J/A+AS/104/365 See VI/96/

J/A+AS/104/379/  refs.dat table1.dat table2.dat  (30Kb)

BATTINELLI P., BRANDIMARTI A., CAPUZZO-DOLCETTA R. Astron. Astrophys. Suppl. Ser. 104, 379 (1994) Integrated photometric properties of open clusters

J/A+AS/104/429/  table4 table6  (477Kb)

OLSEN E.H.

Astron. Astrophys. Suppl. Ser. 104, 429 (1994) A large, complete, volume-limited sample of G-type dwarfs. I. Completion of Stroemgren uvby photometry

J/A+AS/104/509 See VI/82/

J/A+AS/104/529/  table4  (79Kb)

Takata T., Yamada T., Saito M., Chamaraux P., Kazes I. Astron. Astrophys. Suppl. Ser. 104, 529 (1994) Search and redshift survey for IRAS galaxies behind the Northern Milky Way.

J/A+AS/104/557/  table3a table3b  (56Kb)

TERRANEGRA L., CHAVARRIA-K C., DIAZ S., GONZALEZ-PATINO D. Astron. Astrophys. Suppl. Ser. 104, 557 (1994) uvby-beta photometry of Orion population stars and related objects associated with six galactic star-forming regions

J/A+AS/105/15/  table7  (186Kb)

Tian K.P., Zhao J.L., van Leeuwen F. Astron. Astrophys. Suppl. Ser. 105, 15 (1994) Proper motions of stars in the region of open cluster NGC 2286 and its membership study

J/A+AS/105/29 See VI/96/

J/A+AS/105/39 See VI/96/

J/A+AS/105/47/  table1 table2  (32Kb)

ROVITHIS-LIVANIOU H., ROVITHIS P. Astron. Astrophys. Suppl. Ser. 105, 47 (1994) The V505 Sagittarii system: new B,V photoelectric observations and light curve analysis

J/A+AS/105/67/  table1 table2 table3  (11Kb)

Stickel M., Kuehr H. Astron. Astrophys. Suppl. Ser. 105, 67 (1994) Optical Spectra and redshifts of 1 Jy, S4 and S5 radio source identifications.V.

J/A+AS/105/77/ table1.dat table2.dat table3.dat table4.dat table5.dat table6.dat table7.dat table8.dat table9.dat table10a.dat table10b.dat table10c.dat (46Kb) Cornille M., Dubau J., Faucher P., Bely-Dubau F., Blancard C. Astron. Astrophys. Suppl. Ser. 105, 77 (1994) Atomic data and spectral line intensities for O VIII and Fe XVII resonance lines as well as their O VII and Fe XVI satellite lines

J/A+AS/105/125/  refs table1 table1.tex  (94Kb)

Malanushenko V.P., Polosukchina N.S., Weiss W.W. Astron. Astrophys. Suppl. Ser. 105, 125 (1994) Catalogue of CP stars with references to short time scale variability

J/A+AS/105/145/  table1 table2  (60Kb)

PETERSEN J.O. Astron. Astrophys. Suppl. Ser. 105, 145 (1994) The RR Lyrae stars of Omega Centauri Revisited I. Catalogue of Fourier decomposition parameters

J/A+AS/105/179/  table2.dat  (47Kb)

Wroblewski H., Torres C. Astron. Astrophys. Suppl. Ser. 105, 179 (1994) New proper-motion stars south of declination -40deg and right ascension between 16h and 24h

J/A+AS/105/211/  table1 table2 table3  (123Kb)

Stickel M., Meisenheimer K., Kuehr H. Astron. Astrophys. Suppl. Ser. 105, 211 (1994) The optical identification status of the 1 Jy radio source catalogue

J/A+AS/105/243/  table1  (26Kb)

Dimitrijevic M.S., Sahal-Brechot S. Astron. Astrophys. Suppl. Ser. 105, 243 (1994) Stark broadening of Be I spectral lines

J/A+AS/105/245/  table1  (43Kb)

Dimitrijevic M.S., Sahal-Brechot S. Astron. Astrophys. Suppl. Ser. 105, 245 (1994) Stark broadening of spectral lines of multicharged ions of astrophysical interest. XI: Al XI and Si XII

J/A+AS/105/301 See I/208/

J/A+AS/105/305/  table4.dat  (858Kb)

Bohlin R.C., Henrichs H.F., Nichols J.S. Astron. Astrophys. Suppl. Ser. 105, 305 (1994) High-dispersion microfiche atlas of OB stars for the first 10 years of IUE

J/A+AS/105/311/  fig3 fig4 fig6 fig7 table1 table3 table4 table6  (7.8 Mb)

FLUKS M.A., PLEZ B., THE P.S., DE WINTER D., WESTERLUND B.E. STEENMAN H.C. Astron. Astrophys. Suppl. Ser. 105, 311 (1994) On the spectra and photometry of M-giant stars

J/A+AS/105/385/  table8  (13Kb)

KIM K.-T., KRONBERG P.P., DEWDNEY P.E., LANDECKER T.L. Astron. Astrophys. Suppl. Ser. 105, 385 (1994) Radio observations of the Coma cluster of galaxies and its intermediate vicinity. I. Radio data and optical identifications

J/A+AS/105/403/  table2  (84Kb)

Kim K.-T. Astron. Astrophys. Suppl. Ser. 105, 403 (1994) Radio observations of the Coma cluster of galaxies and its intermediate vicinity. II. A master list of radio sources

J/A+AS/105/429/  table1 table2 table3  (14Kb)

CARRASCO G., LOYOLA P. Astron. Astrophys. Suppl. Ser. 105, 429 (1994) Optical positions of the radio stars observed with the Repsold Meridian Circle at Cerro Calan, National Observatory

J/A+AS/105/433/  tables tables.tex  (74Kb)

LONGO G., ZAGGIA S., BUSARELLO G., RICHTER G. Astron. Astrophys. Suppl. Ser. 105, 433 (1994) Kinematics of 14 early-type galaxies

J/A+AS/105/481 See VII/184/

J/A+AS/105/503/  table1 table2  (23Kb)

Balega I.I., Balega Y.Y., Belkin I.N., Maximov A.F., Orlov V.G., Pluzhnik E.A., Shkhagosheva Z.U., Vasyuk V.A. Astron. Astrophys. Suppl. Ser. 105, 503-506 (1994) Binary star speckle measurements during 1989-1993 from the SAO 6m and 1m telescopes in Zelenchuk

J/A+AS/106/1/  table4 table5  (40Kb)

Gregorini L., de Ruiter H.R., Parma P., Sadler E.M., Vettolani G., Ekers R.D. Astron. Astrophys. Suppl. Ser. 106, 1 (1994) Dumbbell galaxies and multiple nuclei in rich clusters: radio data

J/A+AS/106/21/  note ref table1 table1.tex  (433Kb)

RODRIGUEZ E., LOPEZ DE COCA P., ROLLAND A., GARRIDO R., COSTA V. Astron. Astrophys. Suppl. Ser. 106, 21 (1994) Delta Sct stars: a new revised list

J/A+AS/106/29/  table1 table2  (21Kb)

Ghosh T., Gopal-Krishna, Rao A.P. Astron. Astrophys. Suppl. Ser. 106, 29 (1994) Three-year monitoring of a sample of flat-spectrum radio sources at 327 MHz.

J/A+AS/106/79/  table4 table5  (28Kb)

HEYNDERICKX D., HAUG U. Astron. Astrophys. Suppl. Ser. 106, 79 (1994) A new frequency analysis of photometric observations of the beta Cephei star IL Velorum

J/A+AS/106/119 See VIII/49/

J/A+AS/106/141/  table4 table5 table6 table7 table8  (47Kb)

Spoon H.W.W., De Koter A., Sterken C., Lamers H.J.G.L.M., Stahl O. Astron. Astrophys. Suppl. Ser. 106, 141 (1994) Variability of luminous blue variables. I. Intermediate-band photometry

J/A+AS/106/161/  table2 table3  (244Kb)

Bica E., Ortolani S., Barbuy B. Astron. Astrophys. Suppl. Ser. 106, 161 (1994) The metal-rich globular cluster NGC 6356

J/A+AS/106/165/  table1.dat table2.dat table3.dat table4.dat table5.dat  (37Kb)

Aspin C., Sandell G., Russell A.P.G. Astron. Astrophys. Suppl. Ser. 106, 165 (1994) Near-IR imaging photometry of NGC 1333: I. The embedded PMS stellar population.

J/A+AS/106/199/  tablea tableb  (190Kb)

CAON N., CAPACCIOLI M., D’ONOFRIO M. Astron. Astrophys. Suppl. Ser. 106, 199 (1994) ’Global mapping’ B-band photometry of a complete sample of Fornax and Virgo early-type galaxies

J/A+AS/106/257/  table7 table9  (753Kb)

OLSEN E.H. Astron. Astrophys. Suppl. Ser. 106, 257 (1994) Stroemgren photometry of F- and G-type stars brighter than V=9.6. I. uvby photometry

J/A+AS/106/275/ table1 table2 table3 table4 table5 table6 z001.pst z02.pst z0004.pst z004.pst z05.pst z008.pst (7.8 Mb) Bertelli G., Bressan A., Chiosi C., Fagotto F., Nasi E, Astron. Astrophys. Suppl. Ser. 106, 275 (1994) Theoretical isochrones from models with new radiative opacities

J/A+AS/106/303/  table2.dat table3.dat tables.tex  (256Kb)

Neumann M., Reich W., Fuerst E., Brinkmann W., Reich P., Siebert J., Wielebinski R., Truemper J. Astron. Astrophys. Suppl. Ser. 106, 303 (1994) Multifrequency observations of ROSAT selected radio sources

J/A+AS/106/327/  table2 table3  (18Kb)

NOEL F. Astron. Astrophys. Suppl. Ser. 106, 327 (1994) Observations of the Sun during 1993 with the astrolabe of Santiago

J/A+AS/106/333/  table1.dat table2.dat table3.dat table4.dat  (35Kb)

Adelman S.J., Brown B.H., Caliskan H., Reese D.F., Adelman C.J. Astron. Astrophys. Suppl. Ser. 106, 333 (1994) uvby Photometry of the chemically peculiar stars alpha Andromedae, HD 184905, HR 8216, and HR 8434

J/A+AS/106/339/  table3  (13Kb)

VAZQUEZ R.A., BAUME G., FEINSTEIN A., PRADO P. Astron. Astrophys. Suppl. Ser. 106, 339 (1994) Deep photometry in the core of the open cluster NGC 5606

J/A+AS/106/361/  table2 table3  (29Kb)

XIE G.Z., LI K.H., ZHANG Y.H., LIU F.K., FAN J.H., WANG J.C. Astron. Astrophys. Suppl. Ser. 106, 361 (1994) Simultaneous multirange observations and detection of rapid variability of BL Lacertae objects

J/A+AS/106/373/  table1  (29Kb)

DEMICRAN O., DERMAN E., AKALIN A., SELAM S., MUEYESSEROGLU Z. Astron. Astrophys. Suppl. Ser. 106, 373 (1994) UBV photometry of the contact binary AB Andromedae

J/A+AS/106/377/  table1 table2  (102Kb)

Couteau P., Gili R. Astron. Astrophys. Suppl. Ser. 106, 377 (1994) Measurements of double stars made at Nice observatory. New doubles stars (24th series) discovered at Nice with the 50cm refractor.

J/A+AS/106/397/  table6  (26Kb)

KERSCHBAUM F., HRON J. Astron. Astrophys. Suppl. Ser. 106, 397 (1994) Semiregular variables of types SRa and SRb. New JHKL’M-photometry for 200 stars.

J/A+AS/106/441/  table2  (43Kb)

Noel F. Astron. Astrophys. Suppl. Ser. 106, 441 (1994) Second Astrolabe Catalogue of Santiago

J/A+AS/106/451/  params profiles table1 table2 table3 table4 table5 table6 table7  (382Kb)

DE JONG R.S., VAN DER KRUIT P.C. Astron. Astrophys. Suppl. Ser. 106, 451 (1994) Near-infrared and optical broadband surface photometry of 86 face-on disk dominated galaxies. I. Selection,

observations and data reduction.

J/A+AS/106/559/  refs.dat table1.dat  (69Kb)

Tylenda R., Stasinska G., Acker A., Stenholm B. Astron. Astrophys. Suppl. Ser. 106, 559 (1994) A catalogue of HeII 4686 line intensities in Galactic planetary nebulae

J/A+AS/106/573/  table2  (162Kb)

CARRARO G., ORTOLANI S. Astron. Astrophys. Suppl. Ser. 106, 573 (1994) Deep CCD BV photometry of the poorly studied open cluster NGC 4815

J/A+AS/107/9/  table1  (20Kb)

PETROV P.P., SHCHERBAKOV V.A., BERDYUGINA S.V., SHEVCHENKO V.S., GRANKIN K.N., MEL-NIKOV S.Y. Astron. Astrophys. Suppl. Ser. 107, 9 (1994) Photometric and spectroscopic observations of the spotted T Tauri star V410 Tauri

J/A+AS/107/39/  table1 table2 table3 table4 table5 table6 table7  (28Kb)

Claria J.J., Mermilliod J.-C., Piatti A.E., Minnit D. Astron. Astrophys. Suppl. Ser. 107, 39 (1994) Photometric and Coravel observations of stars in the open cluster IC 2714

J/A+AS/107/235/  table1 table2  (59Kb)

Jasinta D.M.D., Soegiartini E. Astron. Astrophys. Suppl. Ser. 107, 235 (1994) Photographic observations of visual double stars

J/A+AS/107/277/  table3 table4 table5 table6 table7 table8 table9  (21Kb)

CUNOW B., WARGAU W.F. Astron. Astrophys. Suppl. Ser. 107, 277 (1994) Photometric CCD sequences in 7 southern Abell clusters

J/A+AS/107/285/  table1 table2  (28Kb)

GAVAZZI G., RANDONE I. Astron. Astrophys. Suppl. Ser. 107, 285 (1994) A CCD survey of galaxies. III. Observations with the Loiano 1.5m telescope

J/A+AS/107/349/  table1  (45Kb)

Dimitrijevic M.S., Sahal-Brechot S. Astron. Astrophys. Suppl. Ser. 107, 349 (1994) Stark broadening of spectral lines of multicharged ions of astrophysical interest X: Ne VIII and Na IX

J/A+AS/107/353/  table4a.dat table4b.dat  (26Kb)

Lopez-Garcia Z., Adelman S.J. Astron. Astrophys. Suppl. Ser. 107, 353 (1994) Elemental abundance studies of CP stars: The silicon star HD 43819 and the CP Star HD 147550

J/A+AS/107/365/  table1 table2 table3 table4 tables.tex  (120Kb)

ALONSO A., ARRIBAS S., MARTINEZ-ROGER C. Astron. Astrophys. Suppl. Ser. 107, 365 (1994) Broad band JHK infrared photometry of an extended sample of late type dwarfs and subdwarfs

J/A+AS/107/385/  table1 table1.tex table4 table4.tex  (54Kb)

TOBIN W., VITON M., SIVAN J.-P. Astron. Astrophys. Suppl. Ser. 107, 385 (1994) The Spacelab-1 Very Wide Field Survey of UV-excess objects. IV. The performance of the instrument in combination with optical photometry as a means of identifying stars with peculiar properties

J/A+AS/107/403/  table1 table2 table3 table4 table5a table5b table6a table6b  (29Kb)

JUZA K., HARMANEC P., BOZIC H., PAVLOVSKI K., ZIZNOVSKY J. TARASOV A.E., HORN J., KOUB-SKY P. Astron. Astrophys. Suppl. Ser. 107, 403 (1994) Correlated long-term light, colour and spectral variations of the Be star Kappa Dra

J/A+AS/107/445/  table1 table1.tex table2 table2.tex  (134Kb)

HASHIMOTO O. Astron. Astrophys. Suppl. Ser. 107, 445 (1994) Circumstellar dust envelopes of oxygen-rich AGB stars

J/A+AS/107/559/  table1  (62Kb)

VEIGA C.H., VIEIRA MARTINS R.

Astron. Astrophys. Suppl. Ser. 107, 559 (1994) Position of Uranian satellites

J/A+AS/108/1/ table1.dat table2.dat table3.dat table4.dat table5.dat table6.dat table7.dat table8.dat table9.dat table10.dat table11.dat table12.dat table13.dat table14.dat tables.tex (47Kb) Butler K., Zeippen C.J. Astron. Astrophys. Suppl. Ser. 108, 1 (1994) Atomic data from the IRON Project. V. Effective collision strengths for transitions in the ground configuration of oxygen-like ions

J/A+AS/108/9/  table1  (115Kb)

BERSIER D., BURKI G., BURNET M. Astron. Astrophys. Suppl. Ser. 108, 9 (1994) Fundamental parameters of Cepheids: I. Photometric data in the Geneva system

J/A+AS/108/25/  table2 table3 table3.tex  (133Kb)

BERSIER D., BURKI G., MAYOR M., DUQUENNOY A. Astron. Astrophys. Suppl. Ser. 108, 25 (1994) Fundamental parameters of Cepheids: II. Radial velocity data

J/A+AS/108/55/  table1  (23Kb)

RIBONI E., PORETTI E., GALLI G. Astron. Astrophys. Suppl. Ser. 108, 55 (1994) The long-term behaviour of the photometric variability of Beta Cassiopeiae

J/A+AS/108/61/  table1 table1.tex table2 table2a.tex table2b.tex  (36Kb)

FEHMERS G.C., DE GRIJP M.H.K., MILEY G.K., KEEL W.C. Astron. Astrophys. Suppl. Ser. 108, 61 (1994) Warm IRAS source from the point source catalog. IV. Extended optical line emission

J/A+AS/108/79/  source source.tex  (378Kb)

Roettgering H.J.A., Lacy M., Miley G.K., Chambers K.C., Saunders R. Astron. Astrophys. Suppl. Ser. 108, 79 (1994) Samples of ultra-steep spectrum radio sources

J/A+AS/108/151/  table1 table2 table3 table4  (342Kb)

Kaluzny J. Astron. Astrophys. Suppl. Ser. 108, 151 (1994) CCD photometry of distant open clusters: I. Berkeley 22, Berkeley 29 and Berkeley 54

J/A+AS/108/287/  table1 table2 table3 table4 table5 table6  (400Kb)

VERNER D.A., BARTHEL P.D., TYTLER D. Astron. Astrophys. Suppl. Ser. 108, 287 (1994) Atomic data for absorption lines from the ground level at wavelengths greater than 228 A

J/A+AS/108/455/  tables tables.tex  (363Kb)

SALGADO M., BONAZZOLA S., GOURGOULHON E., HAENSEL P. Astron. Astrophys. Suppl. Ser. 108, 455 (1994) High precision rotating neutron star models. II. Large sample of neutron stars properties

J/A+AS/108/485/  table1  (42Kb)

VAN DE STEENE G.C., ZIJLSTRA A.A. Astron. Astrophys. Suppl. Ser. 108, 485 (1994) On an alternative statistical distance scale for planetary nebulae. Catalog with statistical distances to planetary nebulae.

J/A+AS/108/561/  table1  (39Kb)

MARKOVA N. Astron. Astrophys. Suppl. Ser. 108, 561 (1994) How did the spectrum of P Cygni look in 1985? Spectral atlas with complete line identifications in the wavelength range from 3550 to 4800A

J/A+AS/108/571/  table  (117Kb)

VALLENARI A., ORTOLANI S., CHIOSI C. Astron. Astrophys. Suppl. Ser. 108, 571 (1994) Photometry of the young cluster NGC 330 in the Small Magellanic Cloud

J/A+AS/108/595/  table table.tex  (20Kb)

CATALANO F.A., LEONE F. Astron. Astrophys. Suppl. Ser. 108, 595 (1994) The light variations of some CP2 SrCrEu J/A+AS/108/603 See III/184/

J/A+AS/108/611/  ca2 ca2k mg2 table5 table6 tables.tex  (346Kb)

LASTENNET E., FREIRE FERRERO R. Astron. Astrophys. Suppl. Ser. 108, 611 (1994) A Ca II and Mg II luminosity/emission-width data base

J/A+AS/108/621/  table  (88Kb)

PELETIER R.F., VALENTIJN E.A., MOORWOOD A.F.M., FREUDLING W. Astron. Astrophys. Suppl. Ser. 108, 621 (1994) The distribution of dust in Sb’s and Sc’s - K-band infrared imaging of a diameter limited sample of 37 galaxies

J/A+AS/108/653/  ngc6440  (109Kb)

ORTOLANI S., BARBUY B., BICA E. Astron. Astrophys. Suppl. Ser. 108, 653 (1994) The low galactic latitude metal-rich globular cluster NGC 6440

J/A+AS/108/669/  table6 tabwcg.tex tabwcl.tex tabwng.tex tabwnl.tex  (135Kb)

NIEDZIELSKI A., ROCHOWICZ K. Astron. Astrophys. Suppl. Ser. 108, 669 (1994) Low resolution IUE spectra of Wolf-Rayet stars.

J/A+AS/109/1/  table6 table7 table8  (80Kb)

RICHTLER T. Astron. Astrophys. Suppl. Ser. 109, 1 (1995) CCD photometry of the globular clusters NGC 6496, NGC 6624, and NGC 6637

J/A+AS/109/9/  tables  (79Kb)

Gottwald M., Parmar A.N., Reynolds A.P, White N.E., Peacock A., Taylor B.G. Astron. Astrophys. Suppl. Ser. 109, 9 (1995) The EXOSAT GSPC iron line catalog.

J/A+AS/109/29/  table2 table3  (50Kb)

HAMBLY N.C., STEELE I.A., HAWKINS M.R.S., JAMESON R.F. Astron. Astrophys. Suppl. Ser. 109, 29 (1995) Very low mass stars in the galactic cluster Praesepe

J/A+AS/109/125/  table1  (130Kb)

VERNER D.A., YAKOVLEV D.G. Astron. Astrophys. Suppl. Ser. 109, 125 (1995) Analytic fits for partial photoionization cross sections

J/A+AS/109/147/  table1  (58Kb)

Brinkmann W., Siebert J., Reich W., Fuerst E., Reich P., Voges W., Truemper J., Wielebinski R. Astron. Astrophys. Suppl. Ser. 109, 147 (1995) The ROSAT AGN content of the 87GB 5GHz survey: bulk properties of previously optically identified sources

J/A+AS/109/171/  table1 table2 table3 table4  (19Kb)

FLYNN C., SOMMER-LARSEN J., CHRISTENSEN P.R., HAWKINS M.R.S. Astron. Astrophys. Suppl. Ser. 109, 171 (1995) A catalog of blue horizontal branch stars in the outer galactic halo

J/A+AS/109/177 See II/199A/

J/A+AS/109/181 See VI/87/

J/A+AS/109/193/  table1 table2  (15Kb)

Berrington K.A. Astron. Astrophys. Suppl. Ser. 109, 193 (1995) Atomic data from the IRON project. VIII. Electron excitation of the 3d^{4} ^{5}D_{J} ground state fine-structure transitions in Ti-like ions V II, Cr III, Mn IV, Fe V, Co VI and Ni VII

J/A+AS/109/201/  table1.dat table2.dat table3.dat table4.dat table5.dat table6.dat  (206Kb)

Garcia J.R., Cebral J.R., Scoccimarro E.R., Wahnon P., Arena R., Bazterra V., Pellizza L., Risi A., Rodriguez M.L., Zimmermann M. Astron. Astrophys. Suppl. Ser. 109, 201 (1995) A catalogue of variable stars in the lower instability strip

J/A+AS/109/263/  table1  (243Kb)

BRETT J.M. Astron. Astrophys. Suppl. Ser. 109, 263 (1995) Opacity sampling model photospheres for M dwarfs. II. The grid

J/A+AS/109/267/  table1.dat table2.dat table3.dat table4.dat table5.dat  (79Kb)

Giommi P., Ansari S.G., Micol A. Astron. Astrophys. Suppl. Ser. 109, 267 (1995) Radio to X-ray energy distribution of BL Lacertae objects

J/A+AS/109/293/  coeff1 coeff2 diffb diffw totb totw  (323Kb)

GREBEL E.K., ROBERTS W.J. Astron. Astrophys. Suppl. Ser. 109, 293 (1995) Heterochromatic extinction. I. Dependence of interstellar extinction on stellar temperature, surface gravity, and metallicity.

J/A+AS/109/313/  coeff1.dat coeff2.dat diffb.dat diffw.dat totb.dat totw.dat  (311Kb)

Roberts W.J., Grebel E.K. Astron. Astrophys. Suppl. Ser. 109, 313 (1995) Heterochromatic extinction. II. Dependence of interstellar extinction on stellar temperature, surface gravity, and metallicity.

J/A+AS/109/329 See II/200/

J/A+AS/109/341/  table5 table6  (89Kb)

Rifatto A., Longo G., Capaccioli M. Astron. Astrophys. Suppl. Ser. 109, 341 (1995) The UV properties of normal galaxies. II. The non-IUE data.

J/A+AS/109/375/  notes.dat refs.dat table1.dat table2.dat  (186Kb)

Ahumada J., Lapasset E. Astron. Astrophys. Suppl. Ser. 109, 375 (1995) Catalogue of blue stragglers in open clusters

J/A+AS/109/425/  table8 table9 table10 table11  (53Kb)

Clausen J.V., Gimnez A., van Houten C.J. Astron. Astrophys. Suppl. Ser. 109, 425 (1995) Four-colour photometry of eclipsing binaries. XXVIIb. A photometric analysis of the (possible) Ap system AO Velorum

J/A+AS/109/441 See VI/118/

J/A+AS/109/447/  photom table1 table1.tex table2 table2.tex  (3.1 Mb)

GRISON P., BEAULIEU J.-P., PRITCHARD J. D., TOBIN W., FERLET R., VIDAL-MADJAR A., GUIB-ERT J., ALARD C., MOREAU O., TAJAHMADY F., MAURICE E., PREVOT L., GRY C., AUBOURG E., BAREYRE P., BREHIN S., GROS M., LACHIEZE-REY M., LAURENT B., LESQUOY E., MAGNEVILLE C., MILSZTAJN A., MOSCOSO L., QUEINNEC F., RENAULT C., RICH J., SPIRO M., VIGROUX L., ZYLBERAJCH S., ANSARI R., CAVALIER F., MONIEZ M. Astron. Astrophys. Suppl. Ser. 109, 447 (1995) EROS catalogue of eclipsing binary stars in the bar of the Large Magellanic Cloud

J/A+AS/109/479/  table4 table5  (36Kb)

Samus N.N., Kravtsov V., Pavlov M., Alcaino G., Liller W. Astron. Astrophys. Suppl. Ser. 109, 479 (1995) A new photographic C-M diagram for NGC 4833

J/A+AS/109/487/  table2  (92Kb)

Samus N., Kravtsov V., Pavlov M., Alcaino G., Liller W. Astron. Astrophys. Suppl. Ser. 109, 487 (1995) A new photographic color-magnitude study of M 22 = NGC 6656

J/A+AS/109/501/  table.dat  (244Kb)

Carrasco G., Loyola P. Astron. Astrophys. Suppl. Ser. 109, 501 (1995) Santiago 91, a right ascension catalogue of 3387 stars (equinox J2000)

J/A+AS/109/537/  table10 table11  (33Kb)

Hopp U., Kuhm B., Thiele U., Birkle K., Elsaesser H. Astron. Astrophys. Suppl. Ser. 109, 537 (1995) A redshift survey for faint galaxies towards voids of galaxies

J/A+AS/109/551/  table1  (52Kb)

Dimitrijevic M.S., Sahal-Brechot S. Astron. Astrophys. Suppl. Ser. 109, 551 (1995) Stark broadening spectral lines of multicharged ions of astrophysical interest. XII. O IV and OV

J/A+AS/110/1/  table10  (25Kb)

Borissova J., Spassova N. Astron. Astrophys. Suppl. Ser. 110, 1 (1995) The CCD photometry of the globular cluster Palomar 1.

J/A+AS/110/19/  table1 table2 table3 table4  (36Kb)

Davoust E., Considere S. Astron. Astrophys. Suppl. Ser. 110, 19 (1995) Kinematical observations of pairs of galaxies.

J/A+AS/110/27/  table1  (32Kb)

Wroblewski H., Torres C. Astron. Astrophys. Suppl. Ser. 110, 27 (1995) New proper motion determination of Luyten catalogue stars (LTT) south of declination -40deg and right ascension between 16h and 24h.

J/A+AS/110/81/  table1 table2  (17Kb)

van der Walt D.J., Gaylard M.J., MacLeod G.C. Astron. Astrophys. Suppl. Ser. 110, 81 (1995) New detections of 5_{1}- 6_{0} A^{+}-methanol masers towards IRAS sources

J/A+AS/110/99/  table.tex table1 table2 table3  (25Kb)

FANG C., DING M.D. Astron. Astrophys. Suppl. Ser. 110, 99 (1995) On the spectral characteristics and atmospheric models of two types of white-light flares

J/A+AS/110/107/  notes table2 table3 table4  (40Kb)

Meyer C., Rabbia Y., Froeschle M., Helmer G., Amieux G. Astron. Astrophys. Suppl. Ser. 110, 107 (1995) Observations of lunar occultations at Observatoire de la Cote d’Azur

J/A+AS/110/131/  table4a table4a.tex table4b table4b.tex table6.tex table6a table6b  (175Kb)

Rampazzo R., Reduzzi L., Sulentic J.W., Madejsky R. Astron. Astrophys. Suppl. Ser. 110, 131 (1995) Gaseous and stellar components in mixed pairs of galaxies. I. The data.

J/A+AS/110/177/  table1 table2  (160Kb)

DUFLOT M., FEHRENBACH C., MANNONE R., BURNAGE R., GENTY V. Astron. Astrophys. Suppl. Ser. 110, 177 (1995) Mesure de vitesses radiales. VII. Accompagnement au sol du programme d’observation du satellite HIPPAR-COS. Radial velocities. VII. Ground based measurements for HIPPARCOS.

J/A+AS/110/269/  table3  (237Kb)

Weinberger R., Saurer W., Seeberger R. Astron. Astrophys. Suppl. Ser. 110, 269 (1995) Penetrating the ”zone of avoidance”. I. A compilation of optically identified extragalactic objects within —b— = 5

J/A+AS/110/279/  tab3.tex tab3cont.tex table3a table3b  (180Kb)

SPERANDIO M., CHINCARINI G., RAMPAZZO R., DE SOUZA R. Astron. Astrophys. Suppl. Ser. 110, 279 (1995) Optical studies of galaxies in cluster. Observation of spirals in Virgo.III

J/A+AS/110/313/  tables.dat  (76Kb)

Flin P., Trevese D., Cirimele G., Hickson P. Astron. Astrophys. Suppl. Ser. 110, 313 (1995) Properties of nearby clusters of galaxies. II. A 151, A 637, A 646, A 649, A 655, A 1132, A 1314, A 1377, A 1570, A 1589

J/A+AS/110/329/  table1 table2 table3  (132Kb)

DIAZ-CORDOVES J., CLARET A., GIMENEZ A. Astron. Astrophys. Suppl. Ser. 110, 329 (1995) Linear and non-linear limb-darkening coefficients for LTE model atmospheres

J/A+AS/110/351/  table1 table1.tex table2 table2.tex  (24Kb)

SANCHEZ M., PARRA F., SOLER M., SOTO R. Astron. Astrophys. Suppl. Ser. 110, 351 (1995) L’astrolabe du ROA. Observations du Soleil en 1992. Observations of the Sun at ROA astrolabe in 1992.

J/A+AS/110/367 See III/182/

J/A+AS/110/371/  table  (26Kb)

Soares D.S.L., de Souza R.E., de Carvalho R.R., Couto da Silva T.C.

Astron. Astrophys. Suppl. Ser. 110, 371 (1995) Southern binary galaxies. I. A sample of isolated pairs.

J/A+AS/110/411/  table2  (33Kb)

BEM J., SZCZODROWSKA-KOZAR B. Astron. Astrophys. Suppl. Ser. 110, 411 (1995) High order f and g power series for orbit determination.

J/A+AS/110/419/  table1  (237Kb)

VISSER A.E., RILEY J.M., ROETTGERING H.J.A , WALDRAM E.M. Astron. Astrophys. Suppl. Ser. 110, 419 (1995) The 7C survey of radio sources at 151 MHz. A 418-square-degree region centred at RA = 17h, Dec = +65deg

J/A+AS/110/469/  table3a table3b table3c  (80Kb)

BADE N., FINK H. H., ENGELS D., VOGES W., HAGEN H.-J., WISOTZKI L., REIMERS D. Astron. Astrophys. Suppl. Ser. 110, 469 (1995) AGN from the ROSAT all-sky survey

J/A+AS/110/533/  tables  (180Kb)

LINDE P., LYNGA G., WESTERLUND B.E. Astron. Astrophys. Suppl. Ser. 110, 533 (1995) A study of clusters and field stars in two regions in the Large Magellanic Cloud. I. CCD photometry in B and V.

J/A+AS/110/553/  table3  (15Kb)

SOKOLOV N.A. Astron. Astrophys. Suppl. Ser. 110, 553 (1995) The determination of Teff of B, A and F main sequence stars from the continuum between 3200 A and 3600 A.

J/A+AS/111/17/  table  (17Kb)

CASTRO S., BARBUY B., BICA E., ORTOLANI S., RENZINI A. Astron. Astrophys. Suppl. Ser. 111, 17 (1995) High-resolution abundance analysis of two stars near the globular cluster NGC 6522 projected in Baade’s window.

J/A+AS/111/41/  gen.dat hd27309.dat hd111133.dat ubvy.dat  (63Kb)

North P., Adelman S.J. Astron. Astrophys. Suppl. Ser. 111, 41 (1995) Stroemgren and Geneva photometry of the magnetic CP stars 56 Tauri, HD 111133, HD 126515 and HD 215441.

J/A+AS/111/57/  table6 table7 table8  (179Kb)

OLOFSSON K. Astron. Astrophys. Suppl. Ser. 111, 57 (1995) Synthetic absorption line strengths in star forming regions, and the effects on the helium abundance determination

J/A+AS/111/75/  table4  (54Kb)

SAGAR R., CANNON R.D. Astron. Astrophys. Suppl. Ser. 111, 75 (1995) A deep UBVRI CCD photometric study of the moderately young southern open star cluster NGC 4755 = κCrucis

J/A+AS/111/85/  table2 table5  (23Kb)

VAZQUEZ R.A., WILL J.-M., PRADO P., FEINSTEIN A. Astron. Astrophys. Suppl. Ser. 111, 85 (1995) A detailed study of the open cluster Pismis 20

J/A+AS/111/95/  table1 table2 table3 table4  (22Kb)

ZINCHENKO I., MATTILA K., TORISEVA M. Astron. Astrophys. Suppl. Ser. 111, 95 (1995) Studies of dense molecular cores in regions of massive star formation II. CSJ=2-1 survey of southern H_{2}O masers in the longitude range l=260deg-310deg.

J/A+AS/111/169/  table1.dat table2.dat table3.dat table7.dat  (152Kb)

Moreau O., Reboul H. Astron. Astrophys. Suppl. Ser. 111, 169 (1995) Systematic search for UV-excess quasar candidates in 40 square degrees at the North Galactic Pole

J/A+AS/111/229/  table1  (16Kb)

ABAD C., DELLA PRUGNA F. Astron. Astrophys. Suppl. Ser. 111, 229 (1995) CCD measurements of visual binaries

J/A+AS/111/237/  table1 table2 table3  (64Kb)

Lewis B.M., David P., Le Squeren A.M. Astron. Astrophys. Suppl. Ser. 111, 237 (1995) Mainline OH detection rates from blue circumstellar shells

J/A+AS/111/265/  table1 table2  (24Kb)

Biviano A., Durret F., Gerbal D., Le Fevre O., Lobo C., Mazure A., Slezak E. Astron. Astrophys. Suppl. Ser. 111, 265 (1995) A catalogue of velocities in the central regions of the Coma cluster

J/A+AS/111/275/  table2 table2.tex  (71Kb)

BARRADO D., BYRNE P.B. Astron. Astrophys. Suppl. Ser. 111, 275 (1995) BV(RI)KC CCD photometry of the core of the young open cluster, NGC 5460

J/A+AS/111/311/  table2 table3 table4 table5 table6 table7  (101Kb)

Filipovic M.D., Haynes R.F., White G.L., Jones P.A., Klein U., Wielebinski R. Astron. Astrophys. Suppl. Ser. 111, 311 (1995) A radio continuum study of the Magellanic Clouds. Part IV: Catalogues of radio sources in the Large Magellanic Cloud at 1.40, 2.45, 4.75, 4.85 and 8.55 GHz

J/A+AS/111/399/  table1 table1.tex  (32Kb)

JAIN S.K., BHATT H.C. Astron. Astrophys. Suppl. Ser. 111, 399 (1995) Study of variability of the polarization in Herbig Ae/Be stars

J/A+AS/111/407/  table1  (43Kb)

PARTHASARATHY M., JAIN S.K. Astron. Astrophys. Suppl. Ser. 111, 407 (1995) Red and reddened stars in the region of Cygnus OB2 (VI Cygni) association.

J/A+AS/111/527/ table8 table9 table10 table11 table12 table13 table14 table15 table16 table17 table18 (95Kb) HOPP U., SCHULTE-LADBECK R.E. Astron. Astrophys. Suppl. Ser. 111, 527 (1995) CCD photometry of 11 resolved dwarf irregular galaxies

J/A+AS/112/35/  tables tables.tex  (264Kb)

GARCIA VARGAS M.L., BRESSAN A., DIAZ A.I. Astron. Astrophys. Suppl. Ser. 112, 35 (1995) Predicted emission lines from giant HII regions ionized by aging star clusters

J/A+AS/112/95/  table  (111Kb)

CORRADI W.J.B., FRANCO G.A.P. Astron. Astrophys. Suppl. Ser. 112, 95 (1995) uvby-beta photometry of 1017 stars earlier than G0 in the Centaurus-Crux-Musca-Chamaeleon direction

J/A+AS/112/99/  table3a table3b table3c  (211Kb)

Leinert C., Vaisaenen P., Mattila K., Lehtinen K. Astron. Astrophys. Suppl. Ser. 112, 99 (1995) Measurements of sky brightness at the Calar Alto Observatory

J/A+AS/112/123/  table2  (61Kb)

VAZ L.P.R., CLAUSEN J.V., GIMENEZ A., HELT B.E., JENSEN K.S. Astron. Astrophys. Suppl. Ser. 112, 123 (1995) Four colour photometry of eclipsing binaries. XXXA. Light curves of LZ Centauri

J/A+AS/112/125/  obs/* table1.dat table2.dat  (245Kb)

Hainaut-Rouelle M.-C., Hainaut O.R., Detal A. Astron. Astrophys. Suppl. Ser. 112, 125 (1995) Lightcurves of selected minor planets

J/A+AS/112/257/  table1 table2  (25Kb)

Gavazzi G., Boselli A., Carrasco L. Astron. Astrophys. Suppl. Ser. 112, 257 (1995) A CCD Survey of Galaxies. IV: Observations with the 2.1m Telescope at San Pedro Martir

J/A+AS/112/291/  table1 table2 table3 table4 tables.tex  (57Kb)

Sinachopoulos D., Cuypers J., Lampens P., Oblak E., van Dessel E.L. Astron. Astrophys. Suppl. Ser. 112, 291 (1995) VRI photometry of wide double stars with A- type primaries.

J/A+AS/112/347/  f287 f295 f351 f534 fsa94 fsgp  (47Kb)

CRISTIANI S., LA FRANCA F., ANDREANI P., GEMMO A., GOLDSCHMIDT P., MILLER L., VIO R., BARBIERI C., BODINI L., IOVINO A., LAZZARIN M., CLOWES R., MACGILLIVRAY H., GOUIFFES CH., LISSANDRINI C., SAVAGE A. [1,2] Reference (19YY) Astron. Astrophys. Suppl. Ser. 112, 347 (1995) The ESO Key-Programme ”A Homogeneous Bright QSO Survey” - I The Methods and the ”Deep” Fields

J/A+AS/112/367/  table2 table3 table4  (1.2 Mb)

WILL J.-M., BOMANS D.J., TUCHOLKE H.-J., DE BOER K.S., GREBEL E.K., RICHTLER T., SEGGEWISS W., VALLENARI A. Astron. Astrophys. Suppl. Ser. 112, 367 (1995) CCD Photometry and Astrometry in Selected Regions C and E in the LMC

J/A+AS/112/407/  table1.dat table2.dat  (26Kb)

Kaluzny J., Kubiak M., Szymanski M., Udalski A., Krzeminski W., Mateo M. Astron. Astrophys. Suppl. Ser. 112, 407 (1995) OGLE Catalogue of Variable Stars in the Sculptor Dwarf Spheroidal Galaxy

J/A+AS/112/429/  table1  (11Kb)

FREUDLING W. Astron. Astrophys. Suppl. Ser. 112, 429 (1995) Neutral Hydrogen Observations of Galaxies in the Hercules Supercluster. III. CGCG Fields at the Edge of the Void

J/A+AS/112/439/  table2  (84Kb)

SAMUS N., IPATOV A., SMIRNOV O., KRAVTSOV V., ALCAINO G., LILLER W., ALVARADO F. Astron. Astrophys. Suppl. Ser. 112, 439 (1995) A CCD BVRI color-magnitude study of the globular cluster NGC 5286

J/A+AS/112/457/  tablea tableb tablec  (27Kb)

GRININ V.P., KOLOTILOV E.A., ROSTOPCHINA A. Astron. Astrophys. Suppl. Ser. 112, 457 (1995) Dust around young stars. Photopolarimetric observations of the T Tauri star BM Andromedae

J/A+AS/112/475 See III/183/

J/A+AS/113/31 See II/202/

J/A+AS/113/35/  table3  (11Kb)

AMRAM P., BOULESTEIX J., MARCELIN M., BALKOWSKI C., CAYATTE V., SULLIVAN W.T.III Astron. Astrophys. Suppl. Ser. 113, 35 (1995) H αvelocity fields and rotation curves of galaxies in clusters. III. Nine galaxies in DC 1842-63

J/A+AS/113/51/  rem1-2 rem3 table1 table2 table3 table4a table4b  (133Kb)

Chen P.S., Gao H., Jorissen A. Astron. Astrophys. Suppl. Ser. 113, 51 (1995) A catalogue of associations between IRAS sources and S stars

J/A+AS/113/71/  papdef.tex paper.tex table1 table1-2.tex table2 table3 table3-4.tex table4  (301Kb)

RAMSAUER J., SOLANKI S.K., BIEMONT E. Astron. Astrophys. Suppl. Ser. 113, 71 (1995) Interesting lines in the infrared solar spectrum. II: Unblended lines between 1.0 and 1.8um

J/A+AS/113/115/  data  (169Kb)

Lagerkvist C.-I, Erikson A., Debehogne H., Festin L., Magnusson P., Mottola S., Oja T., de Angelis G., Belskaya I.N., Dahlgren M., Gonano-Beurer M., Lagerros J., Lumme K., Pohjolainen S. Astron. Astrophys. Suppl. Ser. 113, 115 (1995) Physical studies of asteroids XXIX. Photometry and analysis of 27 asteroids

J/A+AS/113/131/  table2 table3  (51Kb)

Noel F. Astron. Astrophys. Suppl. Ser. 113, 131 (1995) Observations of the Sun during 1994 with the astrolabe of Santiago

J/A+AS/113/151/  table4 table5  (25Kb)

DI NELLA H., GARCIA A.M., GARNIER R., PATUREL G. Astron. Astrophys. Suppl. Ser. 113, 151 (1995) Velocity dispersions for elliptical galaxies I. First set of measurements

J/A+AS/113/237/  table3  (58Kb)

Samain D. Astron. Astrophys. Suppl. Ser. 113, 237 (1995) A High Spectral Resolution Atlas of the Balloon Ultraviolet Spectrum of the Sun: 1950-2000 Angstroems

J/A+AS/113/299/  table6.dat table8.dat  (36Kb)

Minniti D. Astron. Astrophys. Suppl. Ser. 113, 299 (1995) Abundances and velocities for open and globular cluster giants: The data

J/A+AS/113/325/  table1a table1b  (38Kb)

Hilton J., Lahulla J.F. Astron. Astrophys. Suppl. Ser. 113, 325 (1995) Distance measurements of Lynds galactic dark nebulae

J/A+AS/113/419 See I/257B/

J/A+AS/113/441/  table4  (10Kb)

Kerschbaum F. Astron. Astrophys. Suppl. Ser. 113, 441 (1995) Semiregular variables of types SRa and SRb. New JHKL’M-photometry for 44 stars.

J/A+AS/113/451/  table2 table5 table8 table9 table10 table11  (95Kb)

Kalinkov M., Kuneva I. Astron. Astrophys. Suppl. Ser. 113, 451 (1995) Superclusters of galaxies. I. The catalog

J/A+AS/113/473/ table3 table4 table5 table6 table7 table8 table9 table10 table11 table12 table13 table14 table15 table16 table17 table18 table19 table20 table21 table22 table23 table24 table25 table26 (83Kb) Fernandez M. Astron. Astrophys. Suppl. Ser. 113, 473 (1995) Photometric observations of pre-main sequence objects

J/A+AS/113/483/  table3 table4 table5 table6 table7 table8 table9  (71Kb)

CARRILLO D., BURKI G., MAYOR M., BURNET M., LAMPENS P., NICOLET B. Astron. Astrophys. Suppl. Ser. 113, 483 (1995) RR Lyrae stars and Baade-Wesselink method. V. Photometric and radial velocity measurements of BH Peg, RX Eri, SS Leo, UU Vir and X Ari

J/A+AS/113/539/  table2  (51Kb)

Morgan D., Hatzidimitriou D. Astron. Astrophys. Suppl. Ser. 113, 539 (1995) A survey of carbon stars in the Small Magellanic cloud

J/A+AS/113/557/  table1  (32Kb)

Veiga C.H., Vieira Martins R. Astron. Astrophys. Suppl. Ser. 113, 557 (1995) CCD astrometric observations of Uranian satellites

J/A+AS/113/579/  table1 table2  (91Kb)

SCHOENING T. Astron. Astrophys. Suppl. Ser. 113, 579 (1995) Stark broadening of the NV 6f,g,h-7f,g,h,i line complex

J/A+AS/114/1/  table1 table2  (1.8 Mb)

Kaluzny J., Rucinski S.M. Astron. Astrophys. Suppl. Ser. 114, 1 (1995) CCD photometry of distant open clusters: II NGC 6791

J/A+AS/114/21/  table2  (19Kb)

Niklas S., Klein U., Braine J., Wielebinski R. Astron. Astrophys. Suppl. Ser. 114, 21 (1995) A radio continuum survey of Shapley-Ames Galaxies at 2.8 cm. I. Atlas of radio data

J/A+AS/114/79/  polariz.dat table69.dat  (38Kb)

Leroy J.L. Astron. Astrophys. Suppl. Ser. 114, 79 (1995) Linear polarimetry of Ap stars. V. A general catalogue of measurements

J/A+AS/114/109/  table6  (23Kb)

Alcala J.M., Krautter, J., Schmitt J.H.M.M., Covino E., Wichmann R., Mundt R. Astron. Astrophys. Suppl. Ser. 114, 109 (1995)

A study of the Chamaeleon star forming region from the ROSAT all-sky survey. I. X-ray observations and optical identifications

J/A+AS/114/175/  table2a table2b table2c table2d  (4.6 Mb)

Chandra S., Kegel W.H., Le Roy R.J., Hertenstein T. Astron. Astrophys. Suppl. Ser. 114, 175 (1995) Einstein A-coefficients for vibrational transitions in CS

J/A+AS/114/215/  table1 table2 table3 table4 table5 table6 table7  (44Kb)

Magnier E.A., Prins S., van Paradijs J., Lewin W.H.G., Supper R., Hasinger G., Pietsch W., Truemper J. Astron. Astrophys. Suppl. Ser. 114, 215 (1995) New supernova remnant candidates in M31

J/A+AS/114/247/  table1 table2 table3  (96Kb)

Claret A., Diaz-Cordoves J., Gimenez A. Astron. Astrophys. Suppl. Ser. 114, 247 (1995) Linear and non-linear limb-darkening coefficients for the photometric bands R I J H K

J/A+AS/114/253/  table1.dat table3.dat table4.dat table5.dat  (33Kb)

Adelman S.J., Boyce P.W. Astron. Astrophys. Suppl. Ser. 114, 253 (1995) uvby Photometry of the Magnetic Chemically Peculiar Stars HD 11187, HD 15144, 20 Eri, and HR 8933 and the early F Star HD 14940

J/A+AS/114/269 See III/190B/

J/A+AS/114/281/  table2  (10Kb)

Patat F., Carraro G. Astron. Astrophys. Suppl. Ser. 114, 281 (1995) NGC 7762: a forgotten moderate age open cluster

J/A+AS/114/337/  table1 table2  (30Kb)

Jia Guibin, Ma Huiyu, Cen Xuefen, Zhili L., Shengquan L., Guanghua L. Astron. Astrophys. Suppl. Ser. 114, 337 (1995) CCD photometric observations of blazars

J/A+AS/114/341/  table5 table6 table9  (78Kb)

Giovannelli F., Vittone A.A., Rossi C., Errico L., Bisnovatyi-Kogan G.S., Kurt V.G., Lamzin S.A., Larionov M., Sheffer E.K., Sidorenkov V.N. Astron. Astrophys. Suppl. Ser. 114, 341 (1995) Multifrequency monitoring of RU Lupi: I. Observations

J/A+AS/114/363/  table  (137Kb)

Casini R. Astron. Astrophys. Suppl. Ser. 114, 363 (1995) Coefficients for calculating the second-order moments of the polarization profiles of hydrogen lines

J/A+AS/114/465/  table2 table4 table5  (194Kb)

Kreysing H.C., Brunner H., Staubert R. Astron. Astrophys. Suppl. Ser. 114, 465 (1995) A ROSAT XUV Pointed Phase Source Catalogue

J/A+AS/114/487/  table  (71Kb)

Jasinta D.M.D., Raharto M., Soegiartini E. Astron. Astrophys. Suppl. Ser. 114, 487 (1995) Photographic observations of visual double stars

J/A+AS/114/489/  table5  (107Kb)

JORDI C., GALADI-ENRIQUEZ D., TRULLOLS E., LAHULLA F. Astron. Astrophys. Suppl. Ser. 114, 489 (1995) UBVRI-CCD photometry of Cepheus OB3 association

J/A+AS/114/499/  table1  (16Kb)

Markova N., Zamanov R. Astron. Astrophys. Suppl. Ser. 114, 499 (1995) P Cygni - spectral atlas with complete line identifications in the wavelength range from 4840 to 6760 A

J/A+AS/114/527/  table2  (44Kb)

Rifatto A., Longo G., Capaccioli M. Astron. Astrophys. Suppl. Ser. 114, 527 (1995) The UV properties of normal galaxies. III Standard Luminosity Profiles and Total Magnitudes

J/A+AS/114/537/  table3  (21Kb)

Mallik D.C.V., Sagar R., Pati A.K. Astron. Astrophys. Suppl. Ser. 114, 537 (1995) A deep BVI photometric study of the open cluster NGC 2453

J/A+AS/114/549 See VI/118/

J/A+AS/114/565/  table1 table2 tables.tex  (22Kb)

Proust D., Mazure A., Vanderriest C., Sodre L. Jr., Capelato H.V. Astron. Astrophys. Suppl. Ser. 114, 565 (1995) New measurement of radial velocities in clusters of galaxies. IV.

J/A+AS/115/1/  notes table1  (42Kb)

Stickel M., Kuehr H. Astron. Astrophys. Suppl. Ser. 115, 1 (1996) The current optical identification status of the S5 radio source catalogue

J/A+AS/115/11/  table1 table2 table3 table4 table5 table6  (23Kb)

Stickel H., Kuehr M. Astron. Astrophys. Suppl. Ser. 115, 11 (1996) Optical identifications of radio sources from the 1 Jy, S4 and S5 catalogues

J/A+AS/115/41/  table2 table3  (16Kb)

Cutispoto G., Tagliaferri G., Pallavicini R., Pasquini L., Rodono M. Astron. Astrophys. Suppl. Ser. 115, 41 (1996) Photometric and spectroscopic studies of cool stars discovered in EXOSAT X-ray images. III. Photometric properties

J/A+AS/115/59/  table1  (61Kb)

Thorel J.C. Astron. Astrophys. Suppl. Ser. 115, 59 (1996) Mesures et decouvertes d’etoiles doubles effectuees a la lunette de 50 cm de l’Observatoire de Nice. Measures and discoveries of visual double stars made with the 50 cm refractor at the Nice Observatory.

J/A+AS/115/75/  table1 table2  (10Kb)

Costa E., Loyola P. Astron. Astrophys. Suppl. Ser. 115, 75 (1996) Optical astrometry of Benchmark radio sources. III. Positions of 35 southern extragalactic radio sources with respect to the International Reference Stars catalogue.

J/A+AS/115/81/  table1.dat  (72Kb)

Bronfman L., Nyman L.-A., May J. Astron. Astrophys. Suppl. Ser. 115, 81 (1996) A CS(2-1) survey of IRAS point sources with color characteristics of ultra-compact H II regions

J/A+AS/115/97/  refs table1 table2  (110Kb)

Veron-Cetty M.-P., Veron P. Astron. Astrophys. Suppl. Ser. 115, 97 (1996) Astrometry with the Digitized Sky Survey. Positions of 790 AGNs

J/A+AS/115/219/  table2.dat table3.dat  (90Kb)

Pello R., Vilchez-Gomez R. Astron. Astrophys. Suppl. Ser. 115, 219 (1996) Photometric Bgr survey of the distant clusters of galaxies Cl 1613+3104 and Cl 1600+4109

J/A+AS/115/235/  table3  (24Kb)

Reimers D., Koehler T., Wisotzki L. Astron. Astrophys. Suppl. Ser. 115, 235 (1996) The Hamburg/ESO survey for bright QSOs. II. Follow-up spectroscopy of 160 quasars and Seyferts

J/A+AS/115/283/  table.tex table1  (0Kb)

Wu Y., Huang M., He J. Astron. Astrophys. Suppl. Ser. 115, 283 (1996) A catalogue of high velocity molecular outflows

J/A+AS/115/285/ refs.dat table1a.dat table1d.dat table2a.dat table2b.dat table3.dat table4.dat table5.dat table6.dat table7.dat table8.dat table9.dat table10.dat table11.dat table12.dat table13a.dat table13b.dat table14a.dat table14b.dat tables.tex (876Kb) Chan S.J., Henning T., Schreyer K. Astron. Astrophys. Suppl. Ser. 115, 285 (1996) A catalogue of massive young stellar objects

J/A+AS/115/295/  table1  (7Kb)

Hernius O., Lagerkvist C.-I., Lindgren M., Williams G.V., Tancredi G. Astron. Astrophys. Suppl. Ser. 115, 295 (1996) Astrometry of outer Jovian satellites from The Uppsala-ESO Survey of Asteroids and Comets (UESAC)

J/A+AS/115/297/  catalog.dat table4.dat table5.dat  (283Kb)

Lynga G. Astron. Astrophys. Suppl. Ser. 115, 297 (1996) The central field of Omega Centauri

J/A+AS/115/315/  table2 table3  (116Kb)

Clausen J.V., Garcia J.M., Gimenez A., Helt B.E., Jensen K.S., Suso J., Vaz L.P.R. Astron. Astrophys. Suppl. Ser. 115, 315 (1996) Four-colour photometry of eclipsing binaries. XXXVII. Light curves of the B-type system V539 Arae

J/A+AS/115/325/  table1  (7Kb)

Barbon R., Hassan S.M. Astron. Astrophys. Suppl. Ser. 115, 325 (1996) A new study of the young open cluster NGC 7510

J/A+AS/115/339/ table1 table2 table3 table4 table5 table6 table8 table9 table10 table11 table12 table13 table14 table15 table16 table17 (223Kb) Charbonel C., Meynet G., Maeder A., Schaerer D. Astron. Astrophys. Suppl. Ser. 115, 339 (1996) Grid of stellar models. VI. Horizontal branch and early asymptotic giant branch for low mass stars (Z=0.020, 0.001).

J/A+AS/115/345/  table2.dat table3.dat table4.dat table5.dat table6.dat  (80Kb)

Leahy D.A., Roger R.S. Astron. Astrophys. Suppl. Ser. 115, 345 (1996) Radio sources in fields centered near supernova remnant G160.9+2.6 (HB9)

J/A+AS/115/351/  table1  (168Kb)

Dimitrijevic M.S., Sahal-Brechot S. Astron. Astrophys. Suppl. Ser. 115, 351 (1996) Stark broadening spectral lines of multicharged ions of astrophysical interest. XIII. C V and P V

J/A+AS/115/401/  standard table5  (83Kb)

Jordi C., Figueras F., Torra J., Asiain R. Astron. Astrophys. Suppl. Ser. 115, 401 (1996) uvbyHbeta photometry of main sequence A type stars

J/A+AS/115/407/  table1 table2 table3 table4 table5 table6  (43Kb)

Rhee M.-H., Van Albada T.S. Astron. Astrophys. Suppl. Ser. 115, 407 (1996) Short WSRT HI observations of spiral galaxies

J/A+AS/115/439/  table1 table2  (28Kb)

Elfhag T., Booth R.S., Johansson L.E.B., Hoeglund B., Sandqvist A. Astron. Astrophys. Suppl. Ser. 115, 439 (1996) A CO Survey of Galaxies with the SEST and the 20-m Onsala Telescope

J/A+AS/115/469/  table3 table4  (24Kb)

Fu J.-N., Jiang S.-Y., Li Z.-P. Astron. Astrophys. Suppl. Ser. 115, 469 (1996) Photoelectric photometry of 57 and 58 Tauri, two delta Scuti type stars in the Hyades cluster

J/A+AS/115/481/  table2.dat  (44Kb)

Wroblewski H., Torres C. Astron. Astrophys. Suppl. Ser. 115, 481 (1996) New proper-motion stars with declination between -5deg and -30deg and right ascension between 0h and 9h

J/A+AS/116/1/  table1 table2  (23Kb)

Tomov T., Kolev D., Ivanov M., Antov A., Jones A., Mikolajewski M., Lepardo A., Passuello R., Saccavino S., Sostero G., Valentinuzzi T., Bellas-Velidis Y., Dapergolas A., Munari U. Astron. Astrophys. Suppl. Ser. 116, 1 (1996) Monitoring MWC560=V694 Mon in 1990-1995. I. Conventional and high-speed UBV photometry

J/A+AS/116/15/  table1 table1.tex  (21Kb)

Kamphuis J.J., Sijbring D., van Albada T.S. Astron. Astrophys. Suppl. Ser. 116, 15 (1996) Global H I profiles of spiral galaxies

J/A+AS/116/21/  table1 table2 table3 table4 table5  (174Kb)

Andreazza C.M., Vilas-Boas J.W.S. Astron. Astrophys. Suppl. Ser. 116, 21 (1996) Star counts in southern dark clouds: Corona Australis and Lupus

J/A+AS/116/43/  refs.dat table4.dat table4.tex  (94Kb)

Popescu C.C., Hopp U., Hagen H.J., Elsaesser H. Astron. Astrophys. Suppl. Ser. 116, 43 (1996) Search for emission-line galaxies towards nearby voids. Observational data

J/A+AS/116/75/  table2 table2.tex table5 table5.tex  (150Kb)

Vazquez R.A., Baume G., Feinstein A., Prado P. Astron. Astrophys. Suppl. Ser. 116, 75 (1996) Investigation on the region of the open cluster Tr 14

J/A+AS/116/95/  table7 table8 table78a  (130Kb)

Leisy P., Dennefeld M. Astron. Astrophys. Suppl. Ser. 116, 95 (1996) Planetary nebulae in the Magellanic Clouds : I. Carbon abundances and TypeI PNe

J/A+AS/116/157/  table3  (99Kb)

Tornikoski M., Valtaoja E., Terasranta H., Karlamaa K., Lainela M., Nilsson K., Kotilainen J., Laine S., Lahteenmaki A., Knee L.B.G., Botti L.C.L. Astron. Astrophys. Suppl. Ser. 116, 157 (1996) SEST observations of extragalactic radio sources at 3 mm and 1.3 mm

J/A+AS/116/193/  table1 table2  (16Kb)

Casoli F., Dickey J., Kazes I., Boselli A., Gavazzi G., Jore K. Astron. Astrophys. Suppl. Ser. 116, 193 (1996) A 12CO(1-0) survey of spiral galaxies in the region of the Coma supercluster.

J/A+AS/116/203/  table1.dat table2.dat  (74Kb)

Stein P. Astron. Astrophys. Suppl. Ser. 116, 203 (1996) Structure and kinematics of galaxy clusters I. The redshift catalogue

J/A+AS/116/239/ bs788.fit bs1112.fit bs1123.fit bs1790.fit bs2456.fit bs2764.fit bs3170.fit bs3364.fit bs3459.fit bs3903.fit bs4127.fit bs4133.fit bs4174.fit bs4456.fit bs4550.fit bs4590.fit bs4638.fit bs4695a.fit bs4695b.fit bs4883.fit bs4902a.fit bs4902b.fit bs4983.fit bs5020a.fit bs5020b.fit bs5056.fit bs5315.fit bs5645.fit bs5824.fit gl382a.fit gl382b.fit hd93800a.fit hd93800b.fit hd109655.fit hd111180.fit hd112758.fit hd115404.fit hd130871.fit hd157881.fit hd160691.fit hd175674.fit table1 table1.tex table2.tex table3 table3.tex (477Kb) Dallier R., Boisson C., Joly M. Astron. Astrophys. Suppl. Ser. 116, 239 (1996) A library of H band stellar spectra for stellar population analysis purpose

J/A+AS/116/257/  table  (35Kb)

Kaper L., Henrichs H.F., Nichols J.S., Snoek L.C., Volten H., Zwarthoed G.A.A. Astron. Astrophys. Suppl. Ser. 116, 257 (1996) Long- and short-term variability in O-star winds I. Time series of UV spectra for 10 bright O stars.

J/A+AS/116/301 See I/230/

J/A+AS/116/359/  table1 table1.tex table2 table2.tex  (46Kb)

Popovic L.C., Dimitrijevic M.S. Astron. Astrophys. Suppl. Ser. 116, 359 (1996) Stark broadening of Xe II lines.

J/A+AS/116/403/  table3  (10Kb)

Fiorucci M., Tosti G. Astron. Astrophys. Suppl. Ser. 116, 403 (1996) VRI photometry of stars in the fields of 12 BL Lacertae objects

J/A+AS/116/429/  notes table1 table2a table2b table3 table4  (67Kb)

Andreon S., Davoust E., Michard R., Nieto J.-L., Poulain P. Astron. Astrophys. Suppl. Ser. 116, 429 (1996) Morphological classification and structural parameters for early-type galaxies in the Coma cluster.

J/A+AS/116/447/  table1 table2 table3  (18Kb)

Staneva A., Spassova N., Golev V. Astron. Astrophys. Suppl. Ser. 116, 447 (1996) The ellipticities of globular clusters in the Andromeda Galaxy.

J/A+AS/116/473/  table2 table3  (54Kb)

Souchay J., Feissel M., Ma C. Astron. Astrophys. Suppl. Ser. 116, 473 (1996) Precise modeling of nutation and VLBI observations

J/A+AS/116/497/  table2 table5  (45Kb)

Imbert M. Astron. Astrophys. Suppl. Ser. 116, 497 (1996) Cepheids and binarity. II. Radial velocities and spectroscopic orbits of four galactic cepheids: RX Camelopardis, MW Cygni, Z Lacertae and U Vulpeculae

J/A+AS/116/573/  table3  (23Kb)

Quinet P. Astron. Astrophys. Suppl. Ser. 116, 573 (1996) Transition probabilities for forbidden lines of Fe III

J/A+AS/117/1/  table2 table3  (135Kb)

Seeberger R., Saurer W., Weinberger R. Astron. Astrophys. Suppl. Ser. 117, 1 (1996) Penetrating the ”zone of avoidance”. II. Optically detected galaxies in the region 180 = l 240

J/A+AS/117/39/  table2 table2.tex  (66Kb)

Allam S., Assendorp R., Longo G., Braun M., Richter G. Astron. Astrophys. Suppl. Ser. 117, 39 (1996) Far infrared properties of Hickson compact groups of galaxies. I. High resolution IRAS maps and fluxes

J/A+AS/117/93/ hd36395.fit hd38858.fit hd39715.fit hd72184.fit hd72324.fit hd88815.fit hd93800.fit hd112127.fit hd120315.fit hd121370.fit hd123657.fit hd139195.fit hd139669.fit hd141477.fit hd145675.fit hd148856.fit hd156283.fit hd159181.fit hd160762.fit hd161797.fit hd163993.fit hd168532.fit hd176670.fit hd180809.fit hd181984.fit hd185622.fit hd196725.fit hd206936.fit stars.dat (1.1 Mb) Serote Roos M., Boisson C., Joly M. Astron. Astrophys. Suppl. Ser. 117, 93 (1996) Spectra of late type stars from 4800 to 9000 Angstroems

J/A+AS/117/113 See VI/96/

J/A+AS/117/127/  table1 table2 table3  (534Kb)

Dimitrijevic M.S., Sahal-Brechot S. Astron. Astrophys. Suppl. Ser. 117, 127 (1996) Stark broadening of solar Mg I lines

J/A+AS/117/227/  table4.dat table4a.tex table4b.tex  (151Kb)

Alonso A., Arribas S., Martinez-Roger C. Astron. Astrophys. Suppl. Ser. 117, 227 (1996) Determination of effective temperatures for an extended sample of dwarfs and subdwarfs (F0-K5).

J/A+AS/117/261/  table5 table6 table7 table8  (25Kb)

Vennik J., Hopp U., Kovachev B., Kuhn B., Elsaesser H. Astron. Astrophys. Suppl. Ser. 117, 261 (1996) Surface photometry of galaxies in low density regions.

J/A+AS/117/281/  table2 table4 table5  (27Kb)

Jaschek C., Andrillat Y., Jaschek M. Astron. Astrophys. Suppl. Ser. 117, 281 (1996) B[e] stars. I. HD 51585 (= OY Gem)

J/A+AS/117/313 See V/92A/

J/A+AS/117/317/  table1 table2  (18Kb)

Schuster W.J., Nissen P.E., Parrao L., Beers T.C., Overgaard L.P. Astron. Astrophys. Suppl. Ser. 117, 317 (1996) uvby-βphotometry of high-velocity and metal-poor stars. VIII. Stars of very low metal abundance.

J/A+AS/117/335/  table2 table4  (14Kb)

Assafin M., Andrei A.H., Puliaev S.P., Jilinski E.G., Vieira Martins R., Vieira G.G., de Oliveira W.M. Astron. Astrophys. Suppl. Ser. 117, 335 (1996) Optical positions of radio emitting stars.

J/A+AS/117/369/  table2  (201Kb)

Lercher G., Kerber F., Weinberger R. Astron. Astrophys. Suppl. Ser. 117, 369 (1996) Penetrating the ’zone of avoidance’: III. A survey for obscured galaxies in the region 120 l 130,-10 b +10

J/A+AS/117/417/  table1 table2aj table2ak table2b table3  (99Kb)

Heraudeau P., Simien F., Mamon G.A. Astron. Astrophys. Suppl. Ser. 117, 417 (1996) Near-infrared surface photometry of spiral galaxies. I. The data.

J/A+AS/117/467/  catalog  (1.8 Mb)

Garnier R., Paturel G., Petit C., Marthinet M.C., Rousseau J. Astron. Astrophys. Suppl. Ser. 117, 467 (1996) An image database. II. Catalogue between δ=-30deg and δ=70deg.

J/A+AS/117/475/  table2  (21Kb)

Fiorucci M., Tsoti G. Astron. Astrophys. Suppl. Ser. 117, 475 (1996) Automatic optical monitoring of 10 blazars.

J/A+AS/117/519/ table1 table2 table3 table4 table5 table6 table7 table8 table9 table10 table12 table13 (148Kb) Escalera E., MacGillivray H. Astron. Astrophys. Suppl. Ser. 117, 519 (1996) Detection of structures on multiple scales around the South Galactic Pole. A two-dimensional analysis of the COSMOS/UKST Southern Sky Galaxy Catalogue.

J/A+AS/117/557/  table2a table2b table2c table2d  (4.6 Mb)

Chandra S., Maheshwari V.U., Sharma A.K. Astron. Astrophys. Suppl. Ser. 117, 557 (1996) Einstein A-coefficients for vib-rotational transitions in CO

J/A+AS/117/561/ table4a table4b table4c table5a table5b table5c table6a table6b table6c table7a table7b table7c (40Kb) Hure J.M., Roueff E. Astron. Astrophys. Suppl. Ser. 117, 561 (1996) Analytic representations of rovibrational dipole matrix elements for the CO molecule and its isotopomers.

J/A+AS/118/1/  catalog catalog.tex  (678Kb)

Lu L., Manrique W.T., Perdomo R., Podesta R.C., Wang Z., Actis E.L, Zeng F., Zeng Z., Alonso E., Serafino A., Zhao G. , Hormaechea J. Astron. Astrophys. Suppl. Ser. 118, 1 (1996) First catalogue of stars with photoelectric astrolabe in San Juan

J/A+AS/118/7/  table2 table3 table4 table5 table6 table7  (117Kb)

Vitores A.G., Zamorano J., Rego M., Alonso O., Gallego J. Astron. Astrophys. Suppl. Ser. 118, 7 (1996) Photometric and morphological analysis of UCM galaxies. I. Observations and reductions. Morphological classifications

J/A+AS/118/47/  table1 table2 tables.tex  (48Kb)

Chini R., Kruegel E., Lemke R. Astron. Astrophys. Suppl. Ser. 118, 47 (1996) Dust and CO emission in normal spirals. I. The data

J/A+AS/118/65/  table4 table5  (43Kb)

Belloni P., Roeser H.-J. Astron. Astrophys. Suppl. Ser. 118, 65 (1996) Galaxy populations in the distant galaxy clusters. I. Cl0939+472 (z=0.41) and Cl0016+161 (z=0.54). I.

J/A+AS/118/105/  table2  (9Kb)

Ryl’kov V.P., Dement’eva A.A., Alksnis A., Straume J.I. Astron. Astrophys. Suppl. Ser. 118, 105 (1996) Photographic observations of Pluto 1991-1994 with the Baldone Schmidt telescope.

J/A+AS/118/111/  table2 table3 table4 table5  (1.1 Mb)

Heraudeau P., Simien F. Astron. Astrophys. Suppl. Ser. 118, 111 (1996) Optical and I-band surface photometry of spiral galaxies. I. The data

J/A+AS/118/191/  table2 table3 table5 table6 table7  (46Kb)

Juvela M. Astron. Astrophys. Suppl. Ser. 118, 191 (1996) Studies of dense molecular cores in regions of massive star formation. IV. Multitransition CS-study towards southern H_{2}O masers in the longitude range l=308-360deg

J/A+AS/118/231/  table1 table2  (36Kb)

Levato H., Malaroda S., Morrell N., Solivella G., Grosso M. Astron. Astrophys. Suppl. Ser. 118, 238 (1996) Radial velocities and axial rotation of a sample of chemically peculiar stars.

J/A+AS/118/239/  table1  (42Kb)

Hu H., Li D., Li Y., Wang R., Li X. Astron. Astrophys. Suppl. Ser. 118, 239 (1996) The photoelectric astrolabe catalogue of Yunnan Observatory (YPAC)

J/A+AS/118/277/  table7  (130Kb)

Geffert M., Bonnefond P., Maintz G., Guibert J. Astron. Astrophys. Suppl. Ser. 118, 277 (1996) The astrometric accuracy of “Carte du Ciel” plates and proper motions in the field of the open cluster NGC 1647

J/A+AS/118/303/ table3 table4 table5 table6 table7 table8 table9 table10 table11 table12 table13 table14 table15 (341Kb) Kaluzny J., Krzeminski W., Mazur B. Astron. Astrophys. Suppl. Ser. 118 , 330 (1996) Variable stars in the field of the old open cluster NGC 2243

J/A+AS/118/329/  msc.dat notes.dat  (54Kb)

Whiteoak J.B.Z., Green A.J. Astron. Astrophys. Suppl. Ser. 118, 329 (1996) The MOST supernova remnant catalogue (MSC).

J/A+AS/118/381/  table  (17Kb)

Jasinta D.M.D. Astron. Astrophys. Suppl. Ser. 118, 381 (1996) Provisional orbits for three visual binaries

J/A+AS/118/407/  table1 table2 table3 tables.tex  (100Kb)

Nordstroem B., Andersen J., Andersen M.I. Astron. Astrophys. Suppl. Ser. 118, 407 (1996) Critical tests of stellar evolution in open clusters. I. New photometry and radial velocities for NGC 3680

J/A+AS/118/441/  table2 table2.tex table5 table5a.tex table5b.tex  (244Kb)

Schroeder A., Visvanathan N. Astron. Astrophys. Suppl. Ser. 118, 441 (1996) New aperture photometry for 217 galaxies in the Virgo and Fornax clusters.

J/A+AS/118/481/  table2.dat table3.dat  (114Kb)

Berghoefer T.W., Schmitt J.H.M.M., Cassinelli J.P. Astron. Astrophys. Suppl. Ser. 118, 481 (1996) The ROSAT all-sky survey catalogue of optically bright OB-type stars.

J/A+AS/118/495/  table2  (16Kb)

Andrillat Y., Jaschek M., Jaschek C. Astron. Astrophys. Suppl. Ser. 118, 495 (1996) B[e] stars. II. MWC 349 A

J/A+AS/118/503/  table3  (38Kb)

Tian K.P., van Leeuwen F., Zhao J.L., Su C.G. Astron. Astrophys. Suppl. Ser. 118, 503 (1996) Proper motions of stars in the region of the Orion Nebula cluster (C 0532-054)

J/A+AS/118/545/  table4 table4.tex  (29Kb)

Groot P.J., Piters A.J.M., van Paradijs J. Astron. Astrophys. Suppl. Ser. 118, 545 (1996) Rotational velocities of F dwarfs; application of the Fourier-Bessel transformation method

J/A+AS/118/557/  phot tableef tablef tableff tableqf  (191Kb)

De Jong R.S. Astron. Astrophys. Suppl. Ser. 118, 557 (1996) Near-infrared and optical broadband surface photometry of 86 face-on disk dominated galaxies. II. A twodimensional method to determine bulge and disk parameters.

J/A+AS/119/1/  table1 table1.tex table2 table2.tex table3 table3.tex table4 table4.tex  (43Kb)

De Winter D., van den Ancker M.E., Perez M.R., Swings J.-P., The P.S., Johnson S.B., Molster F.J., van Loon J.T.

Astron. Astrophys. Suppl. Ser. 119, 1 (1996) The peculiar B[e] star HD 45677. I. Photometric observations.

J/A+AS/119/37/  table7 table13 table21  (53Kb)

Duijsens M.F.J., van der Hucht K.A., van Genderen A.M., Schwarz H.E., Linders H.P.J., Kolkman O.M. Astron. Astrophys. Suppl. Ser. 119, 37 (1996) Multi-colour photometric and spectroscopic monitoring of the WN5 star EZ Canis Majoris

J/A+AS/119/91 See I/221/

J/A+AS/119/99/  table7 table8  (15Kb)

Quinet P., Le Dourneuf M. Astron. Astrophys. Suppl. Ser. 119, 99 (1996) Transition probabilities for [Ni I] and [Ni II] lines.

J/A+AS/119/111/ high met.dat model4.dat neu neu.dat standard.dat table2a.dat table2b.dat table3a.dat table3b.dat table4a.dat table4b.dat table5a.dat table5b.dat table6a.dat table6b.dat table7a.dat table7b.dat table8a.dat table8b.dat table9a.dat table9b.dat table10a.dat table10b.dat table11a.dat table11b.dat table12a.dat table12b.dat table13a.dat table13b.dat tables.tex (1.3 Mb) Lee H.-H., Bettens R.P.A., Herbst E. Astron. Astrophys. Suppl. Ser. 119, 111 (1996) Fractional abundances of molecules in dense interstellar clouds: a compendium of recent model results

J/A+AS/119/153/ fig3a1.ps fig3a2.ps fig3a3.ps fig3a4.ps fig3a5.ps fig3a6.ps fig3a7.ps fig3a8.ps fig3a9.ps fig3a10.ps fig3a11.ps fig3a12.ps fig3a13.ps fig3a14.ps fig3a15.ps fig3a16.ps fig3a17.ps fig3a18.ps fig3a19.ps fig3a20.ps fig3a21.ps fig3a22.ps fig3a23.ps fig3a24.ps fig3b1.ps fig3b2.ps fig3b3.ps fig3b4.ps fig3b5.ps fig3b6.ps fig3b7.ps fig3b8.ps fig3c1.ps fig3c2.ps fig3c3.ps fig3c4.ps fig3c5.ps fig3c6.ps fig41-3.ps fig44-7.ps fig48-11.ps (2.2 Mb) Mattila K., Vaisaenen P., Appen-Schnur G.F.O.v. Astron. Astrophys. Suppl. Ser. 119, 153 (1996) Sky brightness at the ESO La Silla Observatory 1978 to 1988.

J/A+AS/119/191/  table3  (174Kb)

Samus N., Kravtsov V., Ipatov A., Smirnov O., Alcaino G., Liller W., Alvarado F. Astron. Astrophys. Suppl. Ser. 119, 191 (1996) A CCD BVI color-magnitude study of the metal-rich globular cluster NGC 5927

J/A+AS/119/221/  table1  (52Kb)

Deeg H.J., Ninkov Z. Astron. Astrophys. Suppl. Ser. 119, 221 (1996) Deep CCD photometry and the initial mass function of the core of the OB cluster Berkeley 86.

J/A+AS/119/265/  table3  (130Kb)

Iovino A., Clowes R., Shaver P. Astron. Astrophys. Suppl. Ser. 119, 265 (1996) A large sample of objective prism quasar candidate.

J/A+AS/119/271/  tab5-14 table1 table15 tables.tex  (56Kb)

Fabregat J., Torrejon J.M., Reig P., Bernabeu G., Busquets J., Marco A., Reglero V. Astron. Astrophys. Suppl. Ser. 119, 271 (1996) Be stars in open clusters. I. uvbyβphotometry.

J/A+AS/119/281/  table  (19Kb)

Cutispoto G. Astron. Astrophys. Suppl. Ser. 119, 281 (1996) Long-term monitoring of active stars. V. UBV(RI)c photometry collected in Feb-Mar 1990.

J/A+AS/119/307/  table1  (73Kb)

Le Campion J.-F., Colin J., Geffert M. Astron. Astrophys. Suppl. Ser. 119, 307 (1996) M15, an astrometric standard for CDD observations.

J/A+AS/119/329/  tables  (243Kb)

Slezak E., Infante L., Quintana H. Astron. Astrophys. Suppl. Ser. 119, 329 (1996) CCD sequences in the Shapley supercluster of galaxies. II. The ESO data.

J/A+AS/119/333/  table1 table1.tex  (81Kb)

Macdonald G.H., Gibb A.G., Habing R.J., Millar T.J. Astron. Astrophys. Suppl. Ser. 119, 333 (1996) A 330-360 GHz spectral survey of G 34.3+0.15. I. Data and physical analysis

J/A+AS/119/369/  table1  (77Kb)

Dimitrijevic M.S., Sahal-Brechot S. Astron. Astrophys. Suppl. Ser. 119, 369 (1996) Stark broadening of spectral lines of multicharged ions of astrophysical interest. XIV: Be III and B III.

J/A+AS/119/391/  table3 table4 table5 table6 table7 table8 table9  (260Kb)

Molinari E., Buzzoni A., Chincarini G. Astron. Astrophys. Suppl. Ser. 119, 391 (1996) Photometry of seven clusters of galaxies at intermediate redshift (z=0.15-0.50)

J/A+AS/119/413/  table6  (22Kb)

Galaz G., Ruiz M.T., Thompson I., Roth M. Astron. Astrophys. Suppl. Ser. 119, 413 (1996) NGC 2477: photometry and luminosity function.

J/A+AS/119/459/  refs table2 table2.tex  (44Kb)

Te Lintel Hekkert P., Chapman J.M. Astron. Astrophys. Suppl. Ser. 119, 459 (1996) A search for OH maser emission from post-asymptotic giant branch stars.

J/A+AS/119/483/  table1 table2 tables.tex  (27Kb)

Sinachopoulos D., van Dessel E.L. Astron. Astrophys. Suppl. Ser. 119, 483 (1996) UBV photometry of wide visual double stars. V. Double stars with mainly K- and M- type primaries.

J/A+AS/119/489/  table1 table2 tables.tex  (49Kb)

Mann G., Klassen A., Classen H.-T., Aurass H., Scholz D., MacDowall R. J., Stone R. G. Astron. Astrophys. Suppl. Ser. 119, 489 (1996) Catalogue of solar type II radio bursts observed from September 1990 to December 1993 and their statistical analysis

J/A+AS/119/499/ table1.dat table2.dat table3.dat table4.dat table5.dat table6.dat table7.dat table8.dat table9.dat (19Kb) Sharina M.E., Karachentsev I.D., Tikhonov N.A. Astron. Astrophys. Suppl. Ser. 119, 499 (1996) Photometric distances to NGC 628 and its four companions

J/A+AS/119/509/  aij.dat aij2.dat ls.dat ls2.dat prog.f  (15.7 Mb)

Nahar S.N., Pradhan A.K. Astron. Astrophys. Suppl. Ser. 119, 509 (1996) Atomic data from the IRON Project. XVII. Radiative transition probability for dipole allowed forbidden transitions in Fe III.

J/A+AS/119/523/  levels prog.f ups  (4.9 Mb)

Zhang H.L. Astron. Astrophys. Suppl. Ser. 119, 523 (1996) Atomic data from the Iron Project. XVIII. Electron impact excitation collision strengths and rate coefficients for Fe III.

J/A+AS/119/529/  table1  (67Kb)

Dimitrijevic M.S., Sahal-Brechot S. Astron. Astrophys. Suppl. Ser. 119, 529 (1996) Stark broadening of Sr I spectral lines

J/A+AS/119/547/  table5 table6 table7  (61Kb)

Van der Bliek N.S., Manfroid J., Bouchet P. Astron. Astrophys. Suppl. Ser. 119, 547 (1996) Infrared aperture photometry at ESO (1983-1994) and its future use.

J/A+AS/120/77/  table1 table2  (20Kb)

Filipovic M.D., White G.L., Haynes R.F., Jones P.A., Meinert D., Wielebinski R., Klein U. Astron. Astrophys. Suppl. Ser. 120, 77 (1996) A radio continuum study of the Magellanic Clouds. Part IVa: Catalogues of radio sources in the Large Magellanic Cloud at 2.30 GHz

J/A+AS/120/89/  table1 table2 tables  (295Kb)

Bondi M., Padrielli L., Fanti R., Ficarra A., Gregorini L., Mantovani F. Astron. Astrophys. Suppl. Ser. 120, 89 (1996) A 15 years monitoring program at 408 MHz

J/A+AS/120/99/  table2  (20Kb)

Jaschek M., Andrillat Y., Jaschek C. Astron. Astrophys. Suppl. Ser. 120, 99 (1996) B[e] stars. III. MWC 645

J/A+AS/120/107/  table1.dat  (20Kb)

Veiga C.H., Vieira Martins R. Astron. Astrophys. Suppl. Ser. 120, 107 (1996) CCD observations of Triton

J/A+AS/120/127/  table.dat  (23Kb)

Bouvier J. Astron. Astrophys. Suppl. Ser. 120, 127 (1996) UBVRI photometric monitoring of 7 rapidly rotating late-type dwarfs in the Alpha Persei cluster

J/A+AS/120/139/ table11 table12 table13 table14 table15 table16 table17 table18 table19 table20 table21 table22 table23 table24 table25 table26 table27 table28 table29 table30 table31 table32 table33 table34 table35 table36 table37 table38 table39 table40 table41 table42 table43 table44 table45 table46 table47 table48 table49 table50 table51 table52 table53 table54 table55 table56 (10.7 Mb) Kaluzny J., Kubiak M., Szymanski, M., Udalski A., Lrzeminski W., Mateo M. Astron. Astrophys. Suppl. Ser. 120, 139 (1996) The Optical Gravitational Lensing Experiment. Variable stars in globular clusters - I. Fields 5139A-C in Omega Centauri

J/A+AS/120/153/  table2  (125Kb)

Bica E., Ortolani S., Barbuy B. Astron. Astrophys. Suppl. Ser. 120, 153 (1996) NTT V, I Colour-Magnitude Diagrams of the bulge globular cluster Tonantzintla 2

J/A+AS/120/201/  table1 table2 table3  (17Kb)

Li J., Jin W. Astron. Astrophys. Suppl. Ser. 120, 201 (1996) A compiled catalogue of optical positions of extragalactic radio sources.

J/A+AS/120/257/ tablea1 tablea2 tablea3 tablea4 tablea5 tablea6 tableb1 tableb2 tableb3 tableb4 tableb5 tableb6 (91Kb) Zeilinger W.W.,Pizzella ,A., Amico P., Bertin G., Bertola F., Buson L.M., Danziger I.J., Dejonghe H., Sadler E.M., Saglia R.P., de Zeeuw P.T. Astron. Astrophys. Suppl. Ser. 120, 257 (1996) The distribution of ionized gas in early-type galaxies. II. The velocity field of the ionized gas

J/A+AS/120/267/  table  (92Kb)

Osterloh M., Thommes E., Kania U. Astron. Astrophys. Suppl. Ser. 120, 267 (1996) Detection of periods for T Tauri stars

J/A+AS/120/275/  table3  (15Kb)

Alard C., Terzan A., Guibert J. Astron. Astrophys. Suppl. Ser. 120, 275 (1996) Light curves and periods of Mira variables

J/A+AS/120/283/  table3 table3.tex  (28Kb)

Hofner P., Churchwell E. Astron. Astrophys. Suppl. Ser. 120, 283 (1996) A survey of water maser emission toward ultracompact HII regions

J/A+AS/120/301/ (3Kb) Sanzovo G.C., Singh P.D., Huebner W.F. Astron. Astrophys. Suppl. Ser. 120, 301 (1996) Dust colors, dust release rates, and dust-to-gas ratios in the comae of six comets.

J/A+AS/120/323/  table4 table4.tex table5 table5.tex  (122Kb)

Gallego J., Zamorano J., Rego M., Alonso O., , Vitores A.G. Astron. Astrophys. Suppl. Ser. 120, 323 (1996) Observations of a complete sample of Hαemission-line galaxies. Long-slit spectroscopy of galaxies in UCM lists 1 and 2

J/A+AS/120/357/ table1 table2 table3 table4 table5 table6 table7 table8 table9 table10 table11 table12 table13 table14 table15 tables.dat (37Kb) Infante L., Slezak E., Quintana H. Astron. Astrophys. Suppl. Ser. 120, 357 (1996) CCD Sequences in the Shapley supercluster of Galaxies. I. The LCO-Toronto telescope data.

J/A+AS/120/361/  table4  (35Kb)

Quinet P., Le Dourneuf M., Zeippen C.J. Astron. Astrophys. Suppl. Ser. 120, 361 (1996) Atomic data from the IRON Project. XIX. Radiative transition probabilities for forbidden lines in Fe II

J/A+AS/120/373/  table1-3 tables.tex  (19Kb)

Popovic L.C., Dimitrijevic M.S. Astron. Astrophys. Suppl. Ser. 120, 373 (1996) Stark width for astrophysically important ns-np transitions in Sc II, Y II and Zr II spectra.

J/A+AS/120/429/  table  (52Kb)

Friedemann, C., Reimann, H.-G. and Gurtler, J. Astron. Astrophys. Suppl. Ser. 120, 429 (1996) WW Vulpeculae: photographic magnitudes based on plates of the Harvard Plate Collection

J/A+AS/120/489/  table1 table2  (29Kb)

Gavazzi G., Pierini D., Boselli A., Tuffs R. Astron. Astrophys. Suppl. Ser. 120, 489 (1996) 1.65 micron (H-band) surface photometry of galaxies. I: Observations of 158 galaxies with the Calar Alto 2.2m telescope.

J/A+AS/120/521/  table1.dat table2.dat  (53Kb)

Gavazzi G., Pierini D., Baffa C., Lisi F., Hunt L.K., Randone I., Boselli A. Astron. Astrophys. Suppl. Ser. 120, 521 (1996) 1.65 micron (H-band) surface photometry of galaxies. II: observations of 297 galaxies with the TIRGO 1.5m telescope.

J/A+AS/121/57/  notes.dat refs.txt table1.dat table1.tex  (84Kb)

Catalano F.A., Renson P. Astron. Astrophys. Suppl. Ser. 121, 57 (1997) Fourth supplement to the catalogue of observed periods of Ap stars

J/A+AS/121/59 See VIII/44/

J/A+AS/121/65/  table6 table7 table8 table9  (52Kb)

Harper D., Murray C.D., Beurle K., Williams I.P., Jones D.H.P., Taylor D.B., Greaves S.C. Astron. Astrophys. Suppl. Ser. 121, 65 (1997) CCD astrometry of Saturn’s satellites 1990-1994

J/A+AS/121/71/  appen.tex table4.dat  (26Kb)

Tokovinin A.A. Astron. Astrophys. Suppl. Ser. 121, 71 (1997) Orbits of new spectroscopic components in 7 multiple systems.

J/A+AS/121/95/  table9  (18Kb)

Carreta E., Gratton R.G. Astron. Astrophys. Suppl. Ser. 121, 95 (1997) Abundances for globular cluster giants. I. Homogeneous metallicities for 24 clusters

J/A+AS/121/113/  table2  (109Kb)

Mennickent R.E., Sterken C. Astron. Astrophys. Suppl. Ser. 121, 113 (1997) Near infrared low-resolution spectra of 7 Be stars and AG Carinae

J/A+AS/121/119/ fig9.ps fig11.ps fig13.ps fig20.ps fig23.ps fig26.ps fig27.ps fig33.ps tab6-42 tab43-47 table4 table5 tables.tex (4.3 Mb) Villata M., Raiteri C.M., Ghisellini G., De Francesco G., Bosio S., Latini G., Bucciarelli B., Chiaberge M., Chiumiento G., Cora A., Lanteri L., Lattanzi M.G., Massone G., Peila A., Racioppi F., Smart R.L., Scaltriti F., Anderlucci E. Astron. Astrophys. Suppl. Ser. 121, 119 (1997) Optical photometric monitoring of gamma-ray loud blazars. I. Observations from November 1994 to November 1995.

J/A+AS/121/139/  dipole94.dat heats94.dat prog/* rate95.dat refs.dat  (492Kb)

Millar T.J., Farquhar P.R.A., Willacy K. Astron. Astrophys. Suppl. Ser. 121, 139 (1997) The UMIST database for astrochemistry 1995.

J/A+AS/121/213/  tablea1 tablea2  (21Kb)

Panagi P.M., O’Dell M.A. Astron. Astrophys. Suppl. Ser. 121, 213 (1997)

Spectroscopy of the open cluster Blanco 1. II. Halpha emission as an indicator of relative age

J/A+AS/121/223/  table1 table2 table3 table4 table6 tables.tex  (80Kb)

de Winter D., Koulis C., The P.S., van den Ancker M.E., Perez M.R., Bibo E.A. Astron. Astrophys. Suppl. Ser. 121, 223 (1997) Pre-main sequence candidates in the very young open cluster NGC 6611

J/A+AS/121/243/  table2  (80Kb)

Rybka S.P., Yatsenko A.I. Astron. Astrophys. Suppl. Ser. 121, 243 (1997) GPM1 - a catalog of absolute proper motions of stars with respect to galaxies

J/A+AS/121/247/  notes.dat table2.dat table3.dat  (79Kb)

Grothues H.-G., Oestreicher M.O., Gochermann J., Tappert C., Zaum A., Brugger H.R., Schmidt-Kaler T. Astron. Astrophys. Suppl. Ser. 121, 247 (1997) UBV photometry of galactic foreground and LMC member stars. II. Galactic foreground stars (supplement)

J/A+AS/121/321/  radio table2 table3 table4 table5 table6 table7  (72Kb)

Filipovic M.D., Jones P.A., White G.L., Haynes R.F., Klein U., Wielebinski R. Astron. Astrophys. Suppl. Ser. 121, 321 (1997) A radio continuum study of the Magellanic Clouds. Part V: Catalogues of radio sources in the Small Magellanic Cloud at 1.42, 2.45, 4.75, 4.85 and 8.55 GHz.

J/A+AS/121/327/ appen.tex table5 table5.tex table6 table6.tex table7 table7.tex table8 table8.tex table9 table9.tex table10 table10.tex (58Kb) Bono G., Caputo F., Castellani V., Marconi M. Astron. Astrophys. Suppl. Ser. 121, 327 (1997) Nonlinear investigation of the pulsational properties of RR Lyrae variables

J/A+AS/121/369/  table.dat table3.dat table4.dat tables.tex  (21Kb)

Cutispoto G. Astron. Astrophys. Suppl. Ser. 121, 369 (1997) Long-term monitoring of active stars. VI. UBV(RI)c observations collected in Sep-Oct 1990

J/A+AS/121/407/  table4  (26Kb)

Leisy P., Dennefeld M., Alard C., Guibert J. Astron. Astrophys. Suppl. Ser. 121, 407 (1997) Accurate positions and finding charts of known planetary nebulae in the Large Magellanic Cloud.

J/A+AS/121/451/  table2  (23Kb)

Guarnieri M.D., Carraro G. Astron. Astrophys. Suppl. Ser. 121, 451 (1997) VI photometry of the galactic cluster Berkeley 66

J/A+AS/121/455/  table3 table4 table5  (873Kb)

Kaluzny J. Astron. Astrophys. Suppl. Ser. 121, 455 (1997) CCD Photometry of distant open clusters: III. Berkeley 18.

J/A+AS/121/461/  table1 table10 table11 table12 table13 tables.tex  (66Kb)

Recondo-Gonzalez M.C., Wamsteker W., Clavel J., Rodriguez-Pascual P.M., Vio R., Wang T.-G., Santos-Lleo M., Makino F. Astron. Astrophys. Suppl. Ser. 121, 461 (1997) Ultraviolet variability of the Seyfert 1 galaxy Fairall 9

J/A+AS/121/499/  table6 table7  (14Kb)

Borissova J., Markov H., Spassova N. Astron. Astrophys. Suppl. Ser. 121, 499 (1997) CCD photometry of the globular cluster Palomar 13

J/A+AS/121/507/  table3 table5 table6  (134Kb)

Boselli A., Tuffs R., Gavazzi G., Hippelein H., Pierini D. Astron. Astrophys. Suppl. Ser. 121, 507 (1997) Near infrared surface photometry of late-type Virgo cluster galaxies

J/A+AS/122/1/  tab7-13 tab9-15 table5 table6 template.fit  (9.6 Mb)

Kaluzny J. Astron. Astrophys. Suppl. Ser. 122, 1 (1997) CCD photometry of variable stars in the field of the globular cluster NGC 6397

J/A+AS/122/9/  table6 table7 table8 table9a table9b  (130Kb)

Sagar R., Cannon R.D.

Astron. Astrophys. Suppl. Ser. 122, 9 (1997) Multicolour deep CCD photometric study of the moderately young southern open star clusters NGC 3228, NGC 4103, NGC 5662 and NGC 6087

J/A+AS/122/43/  table1 table2 table3  (23Kb)

Tomov T., Kolev D. Astron. Astrophys. Suppl. Ser. 122, 43 (1997) Monitoring MWC 560 = V694 Mon in 1990-1995. II. Plate spectra

J/A+AS/122/51/ bcdci bcmci calib.doc calib.f kurgen.col ptpgm10i ptpgp00i ptpgp10i xcycm10i xcycp00i xcycp10i (783Kb) Kunzli M., North P., Kurucz R.L., Nicolet B. Astron. Astrophys. Suppl. Ser. 122, 51 (1997) A calibration of Geneva photometry for B to G stars in terms of Teff, logg, and [M/H]

J/A+AS/122/79/ fig2a.eps fig2b.eps fig2c.eps fig2d.eps fig2e.eps fig2f.eps fig2g.eps fig2h.eps fig2i.eps fig2j.eps fig2k.eps table1 table1.tex (6.3 Mb) Sevenster M.N., Chapman J.M., Habing H.J., Killeen N.E.B., Lindqvist M. Astron. Astrophys. Suppl. Ser. 122, 79 (1997) The ATCA/VLA OH 1612 MHz survey. I. Observations of the Galactic Bulge Region

J/A+AS/122/111/  b64.ps b69.ps table2 table3  (252Kb)

Pandey A.K., Durgapal A.K., Bhatt B.C., Mohan V., Mahra H.S. Astron. Astrophys. Suppl. Ser. 122, 111 (1997) Stellar contents of the open clusters Be 64 and Be 69

J/A+AS/122/131/  table3 table4 table6 table7 tables.tex  (41Kb)

Solano E., Fernley J. Astron. Astrophys. Suppl. Ser. 122, 131 (1997) + Erratum vol. 131, p. 195 Spectroscopic survey of delta Scutit stars. I. Rotation velocities and effective temperatures.

J/A+AS/122/163/  table1 table2  (159Kb)

Dimitrijevic M.S., Sahal-Brechot S. Astron. Astrophys. Suppl. Ser. 122, 163 (1997) Stark broadening of Ba I and Ba II spectral lines

J/A+AS/122/201/  supertabular.sty table2 table3 table4 table5 table6 tables.tex  (94Kb)

Motch C., Guillout P., Haberl F., Pakull M.W., Pietsch W., Reinsch K. Astron. Astrophys. Suppl. Ser. 122, 201 (1997) The ROSAT Galactic Plane Survey: analysis of a low latitude sample area in Cygnus. The observations

J/A+AS/122/235/  table2.dat table3.dat table4.dat  (147Kb)

Laurent-Muehleisen S.A., Kollgaard R.I., Ryan P.J., Feigelson E.D., Brinkmann W., Siebert J. Astron. Astrophys. Suppl. Ser. 122, 235 (1997) Radio-loud active galaxies in the northern ROSAT All-Sky Survey. I. Radio identifications

J/A+AS/122/249/  table2 table3 table4 table5 tablea  (47Kb)

Adelman S.J. Astron. Astrophys. Suppl. Ser. 122, 249 (1997) Stroemgren uvby photometry of the magnetic chemically peculiar stars HD 32633, 25 Sex, HR 7224, and HD 200311

J/A+AS/122/271/  table1.dat table4.dat  (121Kb)

Reuter H.-P., Kramer C., Sievers A., Paubert G., Moreno R., Greve A., Leon S., Panis J.F., Ruiz-Moreno M., Ungerechts H., Wild W. Astron. Astrophys. Suppl. Ser. 122, 271 (1997) Millimetre continuum measurements of extragalactic radio sources. IV: Data from 1993-1994.

J/A+AS/122/285/ balmer balmernd brack bracknd lyman lymannd pasch paschnd prog.f size (7.4 Mb) Lemke M. Astron. Astrophys. Suppl. Ser. 122, 285 (1997) Extendend VCS Stark broadening tables for hydrogen-Lyman to Brackett series

J/A+AS/122/399/  ecorr.dat filters.dat kcorr.dat kcorrv.dat sed.dat tables.tex  (335Kb)

Poggianti B.M. Astron. Astrophys. Suppl. Ser. 122, 399 (1997) K and evolutionary corrections from UV to IR

J/A+AS/122/409/  catalog  (852Kb)

Lobo C., Biviano A., Durret F., Gerbal D., Le Fevre O., Mazure A., Slezak E. Astron. Astrophys. Suppl. Ser. 122, 409 (1997)

A photometric catalogue of the Coma cluster core

J/A+AS/122/415/  table2  (17Kb)

de Laverny P., Geoffray H., Jorda L., Kopp M. Astron. Astrophys. Suppl. Ser. 122, 415 (1997) Long-term (UBVRI)c monitoring of 12 southern hemisphere long period variables

J/A+AS/122/447/  table2.dat  (44Kb)

Wroblewski W., Torres C. Astron. Astrophys. Suppl. Ser. 122, 447 (1997) New proper-motion stars with declination between -5deg and-30deg and right ascension between 9h and 13h30m

J/A+AS/122/463/  tab2-14 tab16-18 table15 tables.tex  (122Kb)

Kunkel W.E., Irwin M.J., Demers S. Astron. Astrophys. Suppl. Ser. 122, 463 (1997) Carbon stars in the halo of the Magellanic Clouds: Identification and radial velocity data.

J/A+AS/122/471/ table9 table10 table11 table12 table13 table14 table15 table16 table17 table18 table19 table20 table21 table22 table23 table24 table25 table26 table27 table28 table29 table30 table31 table32 table33 table34 table35 table36 table37 table38 table39 table40 table41 table42 table43 table44 table45 table46 (6.9 Mb) Kaluzny J., Kubiak M., Szymanski, M., Udalski A., Krzeminski W., Mateo M., Stanek K. Astron. Astrophys. Suppl. Ser. 122, 471 (1997) The Optical Gravitational Lensing Experiment. Variable stars in globular clusters - II. Fields 5139D-F in Omega Centauri

J/A+AS/122/483/  table2  (40Kb)

Barbuy B., Ortolani S., Bica E. Astron. Astrophys. Suppl. Ser. 122, 483 (1997) NTT V, I, z photometry of the metal-rich bulge globular cluster Terzan 6

J/A+AS/122/489/  table2.dat table2.tex table3.dat table3.tex  (65Kb)

Guglielmo F., Epchtein N., Arditti F., Sevre F. Astron. Astrophys. Suppl. Ser. 122, 489 (1997) New infrared carbon stars in the IRAS Point Source Catalog

J/A+AS/122/507/  table1.tex table1a table1b  (18Kb)

Hatzidimitriou D., Croke B.F., Morgan D.H., Cannon R.D. Astron. Astrophys. Suppl. Ser. 122, 507 (1997) Kinematics of carbon stars in the outer regions of the Small Magellanic Cloud

J/A+AS/122/515/  table3a table3b table5  (15Kb)

Zhang R.X., Fang M.J., Zhai D.S. Astron. Astrophys. Suppl. Ser. 122, 515 (1997) HL Aurigae: A near-contact binary system

J/A+AS/122/521/  table1.dat table2.dat table3.dat table4.dat  (33Kb)

Simien F., Prugniel P. Astron. Astrophys. Suppl. Ser. 122, 521 (1997) Kinematical data on early-type galaxies. I.

J/A+AS/122/529/ refs.dat table.txt table1.dat table2.dat table3.dat table4.dat table5.dat table6.dat table7.dat (200Kb) Walter H.G., Hering R., de Vegt C. Astron. Astrophys. Suppl. Ser. 122, 529 (1997) Radio stars for linking celestial reference frames

J/A+AS/122/533/  table1  (372Kb)

Dimitrijevic M.S., Sahal-Brechot S. Astron. Astrophys. Suppl. Ser. 122, 533 (1997) Stark broadening of spectral lines of multicharged ions of astrophysical interest. XV: P IV spectral lines.

J/A+AS/122/559/  table3 table4 table5  (22Kb)

Clausen J.V., Larsen S.S., Garcia J.M, Gimenez A., Storm J. Astron. Astrophys. Suppl. Ser. 122, 559 (1997) Secondary standard stars for uvby-beta CCD photometry

J/A+AS/123/1/  fig1.ps table2 table2.tex table3 table3.tex  (154Kb)

Clement R., Garcia M., Reglero V., Clausen J.V., Bravo A., Suso J., Fabregat J., Torrejon J.M., Reig P., Marco A. Astron. Astrophys. Suppl. Ser. 123, 1 (1997)

Four clour photometry of late-type binary systems. I. First uvbyβlight curves of ZZ Ursae Majoris.

J/A+AS/123/5 See III/217/

J/A+AS/123/59/  table2 table2.tex  (139Kb)

Clement R., Garcia M., Reglero V., Clausen J.V., Bravo A., Suso J., Fabregat J. Astron. Astrophys. Suppl. Ser. 123, 59 (1997) Four colour photometry of late-type binary systems. II. New light curves of BH Virginis

J/A+AS/123/63/  table1 table2 table5  (44Kb)

van den Ancker M.E., The P.S., Feinstein A., Vazquez R.A., de Winter D., Perez M.R. Astron. Astrophys. Suppl. Ser. 123, 63 (1997) A multiwavelength study of star formation in the very young open cluster NGC 6530

J/A+AS/123/93/  table1.dat  (14Kb)

Paunzen E., Weiss W.W., Heiter U., North P. Astron. Astrophys. Suppl. Ser. 123, 93 (1997) A consolidated catalogue of lambda Bootis stars

J/A+AS/123/115/  find pg3.ps table1  (8.9 Mb)

Ng Y.K., Schultheis M. Astron. Astrophys. Suppl. Ser. 123, 115 (1997) Field 3 of the Palomar-Groningen Survey. I. Variable stars at the edge of the Sagittarius dwarf galaxy

J/A+AS/123/119/  tablea2.dat  (0Kb)

Einasto M., Tago E., Jaaniste J., Einasto J., Andernach H. Astron. Astrophys. Suppl. Ser. 123, 119 (1997) The supercluster-void network. I. The supercluster catalogue and large-scale distribution

J/A+AS/123/135/  sp/* table1.dat table2.dat  (487Kb)

Ginestet N., Carquillat J.M., Jaschek C., Jaschek M. Astron. Astrophys. Suppl. Ser. 123, 135 (1997) Spectral classifications in the near infrared of stars with composite spectra. II. Study of a sample of 180 stars.

J/A+AS/123/159/  table5 table7  (19Kb)

Galavis M.E., Mendoza C., Zeippen C.J. Astron. Astrophys. Suppl. Ser. 123, 159 (1997) Atomic data from the IRON Project. XXII. Radiative rates for the forbidden transitions within the ground configuration of ions in the Carbon and Oxygen isoelectronic sequences.

J/A+AS/123/209/  table1  (4Kb)

Duemmler R., Ilyin I.V., Tuominen I. Astron. Astrophys. Suppl. Ser. 123, 209 (1997) A new radial velocity curve for the RS Canum Venatoricorum star Sigma Geminorum. Constraints on its physical parameters.

J/A+AS/123/219/  fig1.ps fig2.ps fig3.ps fig4.ps fig6.ps refs table4 table5 table6  (248Kb)

Vigotti M., Vettolani G., Merighi R., Lahulla J., Pedani M. Astron. Astrophys. Suppl. Ser. 123, 219 (1997) The B3-VLA quasar sample

J/A+AS/123/273/  table1 table2  (5.8 Mb)

Hameury J.M., Ritter H. Astron. Astrophys. Suppl. Ser. 123, 273 (1997) Illumination in binaries

J/A+AS/123/305/  tab2-21 tab22-31 tab33-38 table32 tables.tex  (122Kb)

van den Hoek L.B., Groenewegen M.A.T. Astron. Astrophys. Suppl. Ser. 123, 305 (1997) New theoretical yields of intermediate mass stars

J/A+AS/123/329/  table5 table6 table7 table8 table9 table10 table11 table12  (92Kb)

Krautter J., Wichmann R., Schmitt J.H.M.M., Alcala J.M., Neuhaeuser R., Terranegra L. Astron. Astrophys. Suppl. Ser. 123, 329 (1997) New ”weak-line” T Tauri stars in Lupus

J/A+AS/123/353/  table2a table2b table3 table4  (49Kb)

Mathys G., Hubrig S., Landstreet J.D., Lanz T., Manfroid J. Astron. Astrophys. Suppl. Ser. 123, 353 (1997) The mean magnetic field modulus of Ap stars

J/A+AS/123/445/  table2 table3 table4 table5 table6 table7 tablea  (35Kb)

Adelman S.J.

Astron. Astrophys. Suppl. Ser. 123, 445 (1997) On the possible variability of metallic-line and Mercury-Manganese stars

J/A+AS/123/455/  table3 table4  (399Kb)

Will J.-M., Bomans D.J., Dieball A. Astron. Astrophys. Suppl. Ser. 123, 455 (1997) A photometric and spectroscopic study of the association LH 47 in the superbubble N44 in the LMC

J/A+AS/123/473/  sp/* table1.dat  (2.2 Mb)

Montes D., Martin E.L., Fernandez-Figueroa M.J., Cornide M., De Castro E. Astron. Astrophys. Suppl. Ser. 123, 473 (1997) Library of high and mid-resolution spectra in the CaII H + K, Hα, Hβ, NaI D(1), D(2), and HeI D(3) line regions of F, G, K and M field stars

J/A+AS/123/487/  notes.dat table1.dat table1.tex  (55Kb)

Popovic G.M., Pavlovic R. Astron. Astrophys. Suppl. Ser. 123, 487 (1997) CCD measurements of double and multiple stars in Belgrade

J/A+AS/123/507/  table3  (89Kb)

Terzan A., Bernard A., Guibert J. Astron. Astrophys. Suppl. Ser. 123, 507 (1997) A photometric study of the Bright Cloud B in Sagittarius. VIII. 1061 new variable stars.

J/A+AS/123/529/  notes.dat table6.dat  (22Kb)

Gosset E., Moreau O., Surdej J. Astron. Astrophys. Suppl. Ser. 123, 529 (1997) Surveys of ultraviolet-excess quasar candidates in large fields. The area around NGC 450

J/A+AS/123/575/  levels3 levels4 levels5 prog.f table3 table4 table5  (354Kb)

Zhang H.L., Pradhan A.K. Astron. Astrophys. Suppl. Ser. 123, 575 (1997) Atomic data from the Iron Project. XXIII. Relativistic rate coefficients for Fe XXII with inclusion of radiation damping.

J/A+AS/124/1/  table6  (27Kb)

Jerjen H., Dressler A. Astron. Astrophys. Suppl. Ser. 124, 1 (1997) Studies of the Centaurus cluster. I. A catalogue of galaxies in the central region of the Centaurus cluster

J/A+AS/124/13/  table2 table2.tex  (219Kb)

Vazquez R.A., Baume G., Feinstein A., Prado P. Astron. Astrophys. Suppl. Ser. 124, 13 (1997) UBVRI imaging photometry of the open cluster Cr 272

J/A+AS/124/33/  table11 table12 tables.tex  (57Kb)

Shevchenko V.S., Ezhkova O., Tjin A Djie H.R.E., van den Ancker M.E., Blondel P.F.C., de Winter D. Astron. Astrophys. Suppl. Ser. 124, 33 (1997) The FUOR characteristics of the PMS star BN Orionis inferred from new spectroscopic and photometric observations

J/A+AS/124/55/  table1  (100Kb)

Cieslinski D., Jablonski F.J., Steiner J.E. Astron. Astrophys. Suppl. Ser. 124, 55 (1997) Southern and equatorial irregular variables. I. Photoelectric photometry

J/A+AS/124/61/  table5 table5.tex  (286Kb)

Bettoni D., Galletta G. Astron. Astrophys. Suppl. Ser. 124, 61 (1997) A survey of the stellar rotation in barred galaxies

J/A+AS/124/75/  catalog.dat index.dat orbits.dat remarks.dat systems.dat  (367Kb)

Tokovinin A.A. Astron. Astrophys. Suppl. Ser. 124, 75 (1997) MSC - a catalogue of physical multiple stars

J/A+AS/124/85/  table1 table2 tables.tex  (45Kb)

Slysh V.I., Dzura A.M., Val’tts I.E., Gerard E. Astron. Astrophys. Suppl. Ser. 124, 85 (1997) Further search for OH emission from IRAS sources.

J/A+AS/124/123/  fig/* table2  (3.1 Mb)

Massaro E., Feroci M., Matt G. Astron. Astrophys. Suppl. Ser. 124, 123 (1997) An atlas of X- and γ-ray pulse profiles of the Crab pulsar (PSR 0531+21): an analysis of energy and time variations.

J/A+AS/124/143/  table2  (44Kb)

Mennessier M.O., Boughaleb H., Mattei J.A. Astron. Astrophys. Suppl. Ser. 124, 143 (1997) Mean light curves of long-period variables and discrimination between carbon- and oxygen-rich stars.

J/A+AS/124/157/  table2.dat table2.tex  (13Kb)

Geffert M., Klemola A.R., Hiesgen M., Schmoll J. Astron. Astrophys. Suppl. Ser. 124, 157 (1997) Absolute proper motions for the calibration of the Hipparcos proper motion system

J/A+AS/124/205/  table7 table8 table9 table10 table11 table12 tables.tex  (135Kb)

Helmich F.P., van Dishoeck E.F. Astron. Astrophys. Suppl. Ser. 124, 205 (1997) Physical and chemical variations within the W 3 star-forming region. II. The 345 GHz spectral line survey

J/A+AS/124/255/  table1 table2  (366Kb)

Fehrenbach C., Duflot M., Mannone C., Burnage R., Genty V. Astron. Astrophys. Suppl. Ser. 124, 255 (1997) Mesure de vitesses radiales. VIII. Accompagnement au sol du programme d’observation du satellite HIPPAR-COS. Radial velocities. VIII. Ground based measurements for HIPPARCOS.

J/A+AS/124/259/  table6.dat  (1Kb)

Rengelink R.B., Tang Y., de Bruyn A.G., Miley G.K., Bremer M.N., Rottgering H.J.A., Bremer M.A.R. Astron. Astrophys. Suppl. Ser. 124, 259 (1997) The Westerbork Northern Sky Survey (WENSS). I. A 570 square degree Mini-Survey around the North Ecliptic Pole

J/A+AS/124/281/  table1 table2  (21Kb)

Xiang F.Y., Liu Q.Y. Astron. Astrophys. Suppl. Ser. 124, 281 (1997) The variable light curve of BH Virginis

J/A+AS/124/283/  table3 table4 tables.tex  (64Kb)

Pierre M., Oukbir J., Dubreuil D., Soucail G., Sauvageot J.-L., Mellier Y. Astron. Astrophys. Suppl. Ser. 124, 283 (1997) Spectroscopic and photometric study of the bright X-ray cluster A1300

J/A+AS/124/291/  table1 table2 table3 table4 tables.tex  (28Kb)

Niarchos P.G., Hoffmann M., Duerbeck H.W. Astron. Astrophys. Suppl. Ser. 124, 291 (1997) V700 Cygni and AW Virginis: Two W-type W UMa systems with spotted components

J/A+AS/124/299 See III/200B/

J/A+AS/124/353/  table1 table2 table3 tables.tex  (30Kb)

Nakos T., Sinachopoulos D., van Dessel E. Astron. Astrophys. Suppl. Ser. 124, 353 (1997) UBV absolute CCD photometry and differential astrometry of close visual double stars, with G-type primaries.

J/A+AS/124/359/  table1 table2 tables.tex  (47Kb)

Mallik S.V. Astron. Astrophys. Suppl. Ser. 124, 359 (1997) The Ca II triplet lines as diagnostics of luminosity, metallicity and chromospheric activity in cool stars

J/A+AS/124/385/ table1 table5 table6 table7 table8 table9 table10 table11 table12 table13 table14 (32Kb) Zinchenko I., Henning T., Schreyer K. Astron. Astrophys. Suppl. Ser. 124, 385 (1997) Studies of dense molecular cores in regions of massive star formation. V. Structure and kinematics of dense cores from ammonia observations

J/A+AS/124/425/  table1 table2 tables.tex  (111Kb)

Zuccolo R., Selvelli P., Hack M. Astron. Astrophys. Suppl. Ser. 124, 425 (1997) The changes in the UV emission lines of RR Tel from 1978 to 1993. I. The IUE SWP region

J/A+AS/124/437/  catalog  (69Kb)

Rousseau J.M., Perie J.P.

Astron. Astrophys. Suppl. Ser. 124, 437 (1997) Astrometric positions of stars with high proper motions in the Southern Hemisphere

J/A+AS/124/441/  table2  (13Kb)

Andrillat Y., Jaschek C., Jaschek M. Astron. Astrophys. Suppl. Ser. 124, 441 (1997) B[e] stars. IV. HD 45677 = MWC 142

J/A+AS/124/449/  table1.dat table2.dat table3.dat table4.dat tables.tex  (78Kb)

Magazzu A., Martin E.L., Sterzik M.F., Neuhaeuser R., Covino E., Alcala J.M. Astron. Astrophys. Suppl. Ser. 124, 449 (1997) Search for young low-mass stars in a ROSAT selected sample south of the Taurus-Auriga molecular clouds

J/A+AS/124/509/ article-ml.sty fig6a.ps fig6b.ps fig6c.ps fig6d.ps fig6e.ps fig6f.ps fig6g.ps fig6h.ps table1 table1.tex (1.4 Mb) Sevenster M.N., Chapman J.M., Habing H.J., Killeen N.E.B., Lindqvist M. Astron. Astrophys. Suppl. Ser. 124, 509 (1997) The ATCA/VLA OH 1612 MHz survey. II. Observations of the Galactic Disk Region

J/A+AS/124/559/  fig6.ps fig7.ps fig8a.ps fig8b.ps fig9.ps table2 table3 table4 table5 table6  (1.2 Mb)

Karanchentsev I., Drozdovsky I., Kajsin S., Takalo L.O., Heinamaki P. Valtonen M. Astron. Astrophys. Suppl. Ser. 124, 559 (1997) Revised photometric distances to nearby dwarf galaxies in the IC 342/Maffei complex

J/A+AS/125/1/  table2  (133Kb)

Kravtsov V., Ipatov A., Samus N., Smirnov O., Alcaino G., Liller W., Alvarado F. Astron. Astrophys. Suppl. Ser. 125, 1 (1997) NTT CCD photometry of the globular cluster M 79 = NGC 1904 in UBV

J/A+AS/125/11/  cata1 cata2 ei eri.ph hr8714.t7 phoenix vienna  (653Kb)

Strassmeier K.G., Bartus J., Cutispoto G., Rodono M. Astron. Astrophys. Suppl. Ser. 125, 11 (1997) Starspot photometry with robotic telescopes. Continuous UBV and V(RI)c photometry of 23 stars in 1991-1996

J/A+AS/125/65/  table2 table3 table4 table5 tablea  (58Kb)

Adelman S.J. Astron. Astrophys. Suppl. Ser. 125, 65 (1997) uvby Photometry of the magnetic Chemically Peculiar Stars HD 37776, HR 2258, HR 6958, and 108 Aquarii

J/A+AS/125/71/  tables.dat tables.tex  (52Kb)

Cunow B., Duemmler R., Spiekermann G., Ungruhe R., Wargau W.F. Astron. Astrophys. Suppl. Ser. 125, 71 (1997) Photometric CCD sequences for calibration of the ESO/SERC atlas

J/A+AS/125/75/  table2 table3  (11Kb)

Pavlovski K., Harmanec P., Bozic H., Koubsky P., Hadrava P., Kriz S., Ruzic Z., Stefl S. Astron. Astrophys. Suppl. Ser. 125, 75 (1997) UBV photometry of Be stars at Hvar: 1972-1990

J/A+AS/125/149/  table3 table3.tex  (335Kb)

Dere K.P., Landi E., Mason H.E., Monsignori Fossi B.C., Young P.R. Astron. Astrophys. Suppl. Ser. 125, 149 (1997) CHIANTI - an atomic database for emission lines. I. Wavelengths greater than 50 Angstroms

J/A+AS/125/219/  abund newlines tables.tex  (61Kb)

Adelman S.J., Pintado O.I. Astron. Astrophys. Suppl. Ser. 125, 219 (1997) Elemental abundance analyses with Complejo Astronomico El Leoncito REOSC echelle spectrograms. II. μLeporis, 7 Sextantis, HR 4817, and 28 Herculis

J/A+AS/125/229/ lcbm01.cor lcbm01.ori lcbm02.cor lcbm02.ori lcbm03.cor lcbm03.ori lcbm05.cor lcbm05.ori lcbm10.cor lcbm10.ori lcbm15.cor lcbm15.ori lcbm20.cor lcbm20.ori lcbm25.cor lcbm25.ori lcbm30.cor lcbm30.ori lcbm35.cor lcbm35.ori lcbm40.cor lcbm40.ori lcbm45.cor lcbm45.ori lcbm50.cor lcbm50.ori lcbp00.cor lcbp00.ori lcbp01.cor lcbp01.ori lcbp02.cor lcbp02.ori lcbp03.cor lcbp03.ori lcbp05.cor lcbp05.ori lcbp10.cor lcbp10.ori lcbsun.ori lcbvega.ori models.dat read lcb.f (222.7 Mb) Lejeune T., Cuisinier F., Buser R. Astron. Astrophys. Suppl. Ser. 125, 229 (1997) A standard stellar library for evolutionary synthesis. I. Calibration of theoretical spectra.

J/A+AS/125/247/  table1 table2 table3  (31Kb)

Quintana H., Melnick J., Proust D., Infante L. Astron. Astrophys. Suppl. Ser. 125, 247 (1997) Redshift of clusters and galaxies towards the Shapley Concentration

J/A+AS/125/289/  table1  (6Kb)

Zijlstra A.A., Acker A., Walsh J.R. Astron. Astrophys. Suppl. Ser. 125, 289 (1997) Radial velocities of planetary nebulae towards the Galactic bulge

J/A+AS/125/293/  table4  (13Kb)

Pierini D., Gavazzi G., Boselli A., Tuffs R. Astron. Astrophys. Suppl. Ser. 125, 293 (1997) New infrared H and K’ surface photometry of three Tully-Fisher calibrators: N 2366, N 2403 and N4236

J/A+AS/125/313/  appen.dat appen.tex  (26Kb)

Fernley J., Barnes T.G. Astron. Astrophys. Suppl. Ser. 125, 313 (1997) Radial velocities and iron abundances of field RR Lyraes. I.

J/A+AS/125/321/  appen.dat appen.tex  (26Kb)

Solano E., Garrid R., Fernley J., Barnes T.G. Astron. Astrophys. Suppl. Ser. 125, 321 (1997) Radial velocities and iron abundances of field RR Lyraes. II.

J/A+AS/125/329/  table1  (9Kb)

de Souza R.E., de Mello D.F., dos Anjos S. Astron. Astrophys. Suppl. Ser. 125, 329 (1997) Radial velocities and spectrophotometric properties of Southern Binary Galaxies

J/A+AS/125/337/  table2 table3 table4 table5 table6 table7 table8 table9 table10  (93Kb)

Kaluzny J., Krzeminski W., Nalezyty M. Astron. Astrophys. Suppl. Ser. 125, 337 (1997) New variable stars in the globular cluster NGC 288

J/A+AS/125/343/  table1 table2 tables  (1.2 Mb)

Kaluzny J., Kubiak M., Szymanski M., Udalski A., Krzeminski W., Mateo M. Astron. Astrophys. Suppl. Ser. 125, 343 (1997) The Optical Gravitational Lensing Experiment. Variable stars in globular clusters. III. RR Lyrae stars and Pop. II Cepheids in omega Centauri.

J/A+AS/125/399/ fig3 1.ps fig3 2.ps fig3 3.ps fig3 4.ps fig3 5.ps fig3 6.ps fig3 7.ps fig3 8.ps fig3 9.ps fig3 10.ps fig3 11.ps fig3 12.ps fig3 13.ps fig3 14.ps fig3 15.ps fig3 16.ps fig3 17.ps fig3 18.ps fig3 19.ps fig3 20.ps table2.tex table2a table2b (54.7 Mb) Arlot J.E., Ruatti C., Thuillot W., Arsenijevic J., Baptista R., Barroso Jr. J., Bauer C., Berthier J., Blanco C., Bouchet P., Bourgeois J., Bulder H.J.J., Burchi R., Cano J.A., Casas R., Chauvet F., Chis D., Colas F., Colin J., D’Ambrosio V., De Angelis G., De Benedetto G., Denzau H., Desbats J.M., Descamps P., Dipaolantonio A., Dumitrescu A., Farcas L., Federspiel M., Flatres T., Froeschle M., Gherega O., Gomez-Forrellad J.M., Guarro J., Hainaut O., Horvat A., Helmer G., Hube D., Ito Y., Kidger M., Lecacheux J., Le Campion J.F., Le Floch J.C., Mallama A., Martin B.E., Mellilo J.F., Meyer C., Molau S., Montignac G., Morando B., Nicolet B., Nitschelm B., Oprescu G., Piersimoni A., Przewozny D., Protitch-Benishek V., Rapaport M., Riou A., Sacre J.J., Sevre F., Shkodrov V., Souchay J., Takami H., Taylor R., Tholen D.J., Turcu V., Vasundhara R., Vidal J., Vu D.T., White G., Wilds R.P. Astron. Astrophys. Suppl. Ser. 125, 399 (1997) A catalogue of the observations of the mutual phenomena of the Galilean satellites made in 1991 during the PHEMU91 Campaign

J/A+AS/125/407/  table2 table5  (73Kb)

Akalin A., Derman E. Astron. Astrophys. Suppl. Ser. 125, 407 (1997) UBV photometry and period variations of V839 Ophiuchi

J/A+AS/125/419/  table1 table2 table3 table4 table5 table6 table7 table8  (85Kb)

Loup C., Zijlstra A.A., Waters L.B.F.M., Groenewegen M.A.T. Astron. Astrophys. Suppl. Ser. 125, 419 (1997) Obscured AGB stars in the Magellanic Clouds. I. IRAS candidates

J/A+AS/125/439 See VI/118/

J/A+AS/125/459/  tables.dat tables.ps tables.tex  (1.2 Mb)

Trevese D., Cirimele G., Cenci A., Appodia B., Flin P., Hickson P. Astron. Astrophys. Suppl. Ser. 125, 459 (1997)

Properties of nearby clusters of galaxies. III: A 76, A 157, A 407, A 505, A 671, A 779, A 1700, A 2028, A 2040, A 2052, A 2063, A 2065, A 2593, A 2657, A 2670

J/A+AS/125/471/  table2  (74Kb)

Vaz L.P.R., Andersen J., Clausen J.V., Helt B.E., Garcia J.M., Gimenez A., Alencar S.H.P. Astron. Astrophys. Suppl. Ser. 125, 471 (1997) Four-colour photometry of eclipsing binaries. XXXVIII. Light curves of the triple system V906 Scorpii

J/A+AS/125/475/  fig1.ps table1 table2 tables.tex  (609Kb)

Qian S., Liu Q., Yang Y., Gu S., Huang Z. Astron. Astrophys. Suppl. Ser. 125, 475 (1997) Photometric study of the close binary system DD Monocerotis

J/A+AS/125/497/  table2 table2b table3 table3b tablea  (83Kb)

Adelman S.J. Astron. Astrophys. Suppl. Ser. 125, 497 (1997) On the possible variability of the main sequence A stars theta Virginis and 109 Virginis

J/A+AS/125/523 See I/237/

J/A+AS/125/539/  table2 table4  (41Kb)

Uylings P.H.M., Raassen A.J.J. Astron. Astrophys. Suppl. Ser. 125, 539 (1997) Atomic data for interpretation of HgMn spectra: transition probabilities of Mn III

J/A+AS/125/543/  table5 table5.tex  (69Kb)

Ramsbottom C.A., Bell K.L. Astron. Astrophys. Suppl. Ser. 125, 543 (1997) Effective collision strengths for electron-impact excitation of nitrogen-like Mg VI

J/A+AS/126/1/  table1 table2  (58Kb)

Gili R., Couteau P. Astron. Astrophys. Suppl. Ser. 126, 1 (1997) Measurements of double stars made at Nice Observatory with the 74 and 50 cm refractors

J/A+AS/126/3/ fig3p1.ps fig3p2.ps fig3p3.ps fig3p4.ps fig3p5.ps fig3p6.ps fig3p7.ps fig3p8.ps fig3p9.ps fig3p10.ps fig3p11.ps fig3p12.ps table1.dat table3.dat (5.5 Mb) Horellou C., Booth R. Astron. Astrophys. Suppl. Ser. 126, 3 (1997) CO and HI in a southern sample of interacting galaxies. I. The data

J/A+AS/126/15/  table1 table2 table3 table4  (43Kb)

Simien F., Prugniel P. Astron. Astrophys. Suppl. Ser. 126, 15 (1997) Kinematical data on early-type galaxies. II.

J/A+AS/126/21/  table1 table2 table4 table5 table6  (366Kb)

Nordstroem B., Stefanik R.P., Latham D.W., Andersen J. Astron. Astrophys. Suppl. Ser. 126, 21 (1997) Radial velocities, rotations, and duplicity of a sample of early F-type dwarfs

J/A+AS/126/39/  tables tables.ps tables.tex  (112Kb)

Steffen M., Szczerba R., Men’shchikov A., Schoenberner D. Astron. Astrophys. Suppl. Ser. 126, 39 (1997) Hydrodynamical models and synthetic spectra of circumstellar dust shells around AGB stars. I. Stationary solutions.

J/A+AS/126/67/  comafig.ps persfig.ps table1 table2 table3 table4 table5 tables.tex  (148Kb)

Andreon S., Davoust E., Poulain P. Astron. Astrophys. Suppl. Ser. 126, 67 (1997) Morphological classification and structural parameter of galaxies in the Coma and Perseus clusters

J/A+AS/126/73/  catalog.dat  (25Kb)

Lahulla J.F., Aguirre A., Hilton J. Astron. Astrophys. Suppl. Ser. 126, 73 (1997) BVRI photometric sequences for nine selected dark globules

J/A+AS/126/81/  table3.dat table3.tex  (13Kb)

Foster D.C., Byrne P.B., Hawley S.L., Rolleston W.R.J. Astron. Astrophys. Suppl. Ser. 126, 81 (1997) CCD photometry of late-type stars in the young open cluster IC 2602

J/A+AS/126/247/  table1.dat  (108Kb)

Saurer W., Seeberger R., Weinberger R. Astron. Astrophys. Suppl. Ser. 126, 247 (1997) Penetrating the zone of avoidance. IV. An optical survey for galaxies in the region 130 l 180, -5 b +5

J/A+AS/126/267/  dat/* list.dat  (102.3 Mb)

Chavez M., Malagnini M. L., Morossi C. Astron. Astrophys. Suppl. Ser. 126, 267 (1997) An Atlas of high resolution synthetic spectra in the wavelength region 4850-5400 A

J/A+AS/126/273/  notes.dat table1.dat table2.dat  (39Kb)

Muller P. Astron. Astrophys. Suppl. Ser. 126, 273 (1997) Mesures d’etoiles doubles faites a Nice (5eme serie). Double star measurements made at Nice.

J/A+AS/126/281/  fig4.ps table1.dat table1.tex  (205Kb)

Curdt W., Feldman U., Laming J.M., Wilhelm K., Schuehle U., Lemaire P. Astron. Astrophys. Suppl. Ser. 126, 281 (1997) The solar disk spectrum between 660 and 1175 A. (First Order) obtained by SUMER on SOHO

J/A+AS/126/297/  pos2000 table1 table2a table2b table5 table6  (80Kb)

Ratag M.A., Pottach S.R., Dennefeld M., Menzies J. Astron. Astrophys. Suppl. Ser. 126, 297 (1997) Abundances in planetary nebulae near the galactic centre. I. Abundance determinations

J/A+AS/126/311/  table1 table2  (28Kb)

Carrasco G., Ledoux C., Loyola P. Astron. Astrophys. Suppl. Ser. 126, 311 (1997) UBVRI photometry of the FK5 Extension Catalogue Stars.

J/A+AS/126/319/  table2 table3 table4  (99Kb)

Ortolani S., Bica E., Barbuy, B. Astron. Astrophys. Suppl. Ser. 126, 319 (1997) V, I and Gunn z photometry of faint bulge globular clusters: Terzan 10, ESO456-SC38 and UKS 1

J/A+AS/126/325/  table3 table4 tables.tex  (37Kb)

Marx M., Dickey J.M., Mebold U. Astron. Astrophys. Suppl. Ser. 126, 325 (1997) A Catalogue of Compact Radio Sources in and behind the Large Magellanic Cloud

J/A+AS/126/357/  fig4.ps  (116Kb)

Rolleston W.R.J., Byrne P.B. Astron. Astrophys. Suppl. Ser. 126, 357 (1997) Photometrically determined membership of the young, open cluster IC 2391

J/A+AS/126/373/  levels3 prog.f table3  (1.8 Mb)

Zhang H.L., Pradhan A.K. Astron. Astrophys. Suppl. Ser. 126, 373 (1997) Atomic data from the Iron Project. XXVII. Electron impact excitation collision strengths and rate coefficients for Fe IV

J/A+AS/126/401/  figure5.ps figure6.ps figure7.ps figure8.ps figure9.ps figure10.ps table1  (993Kb)

Magnier E.A., Augusteijn T., Prins S., van Paradijs J., Lewin W.H.G. Astron. Astrophys. Suppl. Ser. 126, 401 (1997) Cepheids as tracers of star formation in M 31. I. Observations and identifications.

J/A+AS/126/413/  table1  (189Kb)

Reich P., Reich W., Fuerst E. Astron. Astrophys. Suppl. Ser. 126, 413 (1997) The Effelsberg 21cm radio continuum survey of the Galactic Plane between l=95.5 degree and l=240 degree

J/A+AS/126/471/  ref1 table1 table2  (18Kb)

Tamazian V.S., Theureau G., Coudreau-Durand N. Astron. Astrophys. Suppl. Ser. 126, 471 (1997) 21-cm line observations of galaxies from Kazarian’s lists

J/A+AS/126/479/ notes.tex refs.dat stars.dat table3.dat table4.dat table6.dat table6.tex tables.ps (626Kb) Garcia-Lario P., Manchado A., Pych W., Pottasch S.R. Astron. Astrophys. Suppl. Ser. 126, 479 (1997) Near infrared photometry of IRAS sources with colours like planetary nebulae. III.

J/A+AS/126/503/  tablea1.dat tablea1.tex  (110Kb)

Ma J., Peng Q., Chen R., Ji Z., Tu C.

Astron. Astrophys. Suppl. Ser. 126, 503 (1997) Scaleheights of 486 southern spiral galaxies and some statistical correlation

J/A+AS/126/509/  table4 table5 table6  (76Kb)

Molthagen K., Wendker H.J., Briel U.G. Astron. Astrophys. Suppl. Ser. 126, 509 (1997) Multiwavelength observations of the field HS 47.5/22 in Ursa Major. I. The X-ray catalogue of a medium deep ROSAT survey.

J/A+AS/126/519/  table1 table2 table3 table4  (37Kb)

Simien F., Prugniel P. Astron. Astrophys. Suppl. Ser. 126, 519 (1997) Kinematical data on early-type galaxies. III.

J/A+AS/127/1/  map0060.fit map0100.fit refs  (2.1 Mb)

Leinert C., Bowyer S., Haikala L.K., Hanner M.S., Hauser M.G., Levasseur-Regourd A.-C., Mann I., Mattila K., Reach W.T., Schlosser W., Staude H.J., Toller G. N., Weiland J.L., Weinberg J.L., Witt A.N. Astron. Astrophys. Suppl. Ser. 127, 1 (1998) The 1997 reference of diffuse night sky brightness

J/A+AS/127/101/  table2  (119Kb)

Seeberger R., Saurer W. Astron. Astrophys. Suppl. Ser. 127, 101 (1998) Penetrating the ”zone of avoidance” V. An optical survey for hidden galaxies in the region 90 =l =110, -10 =b =+10

J/A+AS/127/117/  table1 table2 table3  (467Kb)

Prugniel P., Zasov A., Busarello G., Simien F. Astron. Astrophys. Suppl. Ser. 127, 117 (1998) A catalogue of spatially resolved kinematics of galaxies: Bibliography

J/A+AS/127/119/  refs.dat table2.dat table2b.dat table3.dat table3b.dat tables.ps tables.tex  (131Kb)

Filipovic M.D., Pietsch W., Haynes R.F., White G.L., Jones P.A., Wielebinski R., Klein U., Dennerl K., Kahabka P., Lazendic J.S. Astron. Astrophys. Suppl. Ser. 127, 119 (1998) A radio continuum study of the Magellanic Clouds. VI. Discrete sources common to radio and X-ray surveys of the Magellanic Clouds.

J/A+AS/127/145/  optical1.dat optical2.dat rassid1.dat rassid2.dat sky.ps  (4Kb)

Bade N., Engels D., Voges W., Beckmann V., Boller T., Cordis L., Dahlem M., Englhauser J., Molthagen K., Nass P., Studt J., Reimers D. Astron. Astrophys. Suppl. Ser. 127, 145 (1998) The Hamburg/RASS Catalogue of optical identifications.

J/A+AS/127/167/  table6  (1.3 Mb)

Richtler T., Grebel E.K., Subramaniam A., Sagar S. Astron. Astrophys. Suppl. Ser. 127, 167 (1998) V,I CCD photometry of metal-rich globular clusters: NGC 6528

J/A+AS/127/181/ fig1-1.ps fig1-2.ps fig1-3.ps fig1-4.ps fig1-5.ps fig1-6.ps fig1-7.ps fig1-8.ps fig1-9.ps fig1-10.ps table1 table1.tex (840Kb) Han F., Mao R.Q., Lu J., Wu Y.F., Sun J., Wang J.S., Pei C.C., Fan Y., Tang G.S., Ji H.R. Astron. Astrophys. Suppl. Ser. 127, 181 (1998) New detections of H2O maser sources on the 13.7m radio telescope of Purple Mountain Observatory

J/A+AS/127/185/  refs.dat table1 table2 tables.tex  (60Kb)

Nyman L.-A., Hall P.J., Olofsson H. Astron. Astrophys. Suppl. Ser. 127, 185 (1998) SiO masers in OH/IR stars, proto-planetary and planetary nebulae

J/A+AS/127/201/ 008717.ps 188309.ps 191439.ps 192674.ps 192687.ps 194750.ps 195112.ps 200623.ps 207259.ps table4 (410Kb) Manfroid J., Burnet M., Renson P. Astron. Astrophys. Suppl. Ser. 127, 201 (1998) Variations of Ap stars in the Geneva system

J/A+AS/127/207/  appen.dat table1.dat table3.dat table4.dat tables.tex  (39Kb)

Cutispoto G. Astron. Astrophys. Suppl. Ser. 127, 207 (1998) Long-term monitoring of active stars. VII. UBV(RI)_{c}photometry collected in March 1991

J/A+AS/127/217/  table table.tex  (72Kb)

D’Odorico V., Cristiani S., D’Odorico S., Fontana A., Giallongo E. Astron. Astrophys. Suppl. Ser. 127, 217 (1998) The absorption spectrum of the QSO PKS 2126-158 (z_{em}=3.27) at high resolution

J/A+AS/127/251/  table2  (53Kb)

Huensch M., Schmitt H.H.M.M., Voges W. Astron. Astrophys. Suppl. Ser. 127, 251 (1998) The ROSAT all-sky survey catalogue of optically bright late-type giants and supergiants

J/A+AS/127/277/  table2 table3 table6 tables.tex  (56Kb)

Kuenzli M., North P. Astron. Astrophys. Suppl. Ser. 127, 277 (1998) Are metallic A-F giants evolved Am stars? Rotation and rate of binaries among giant F stars

J/A+AS/127/295/  table1 table2 tables.tex  (58Kb)

Popovic L.C., Dimitrijevic M.S. Astron. Astrophys. Suppl. Ser. 127, 295 (1998) Stark broadening parameters for Kr II lines from 5s-5p transitions

J/A+AS/127/367/ bj4882.dat bj9229.dat br2936.dat bu9362.dat catalog.dat sumcat.dat sumcat1.ps sumcat1.tex sumcat2.ps sumcat2.tex sumcat3.ps sumcat3.tex sumcat4.ps sumcat4.tex (2.1 Mb) Young C.K., Currie M.J. Astron. Astrophys. Suppl. Ser. 127, 367 (1998) The ”Virgo photometry catalogue”; a catalogue of 1180 galaxies in the direction of the Virgo Cluster’s core

J/A+AS/127/409/  table1 table2  (19Kb)

Karachentseva V.E., Karachentsev I.D. Astron. Astrophys. Suppl. Ser. 127, 409 (1998) A list of new nearby dwarf galaxy candidates

J/A+AS/127/421 See III/199A/

J/A+AS/127/423/  table  (18Kb)

Piatti A.E., Bica E., Claria J.J. Astron. Astrophys. Suppl. Ser. 127, 423 (1998) Fundamental parameters of the highly reddened young open clusters Westerlund 1 and 2

J/A+AS/127/445/  tab2-21 table  (54Kb)

Raiteri C.M., Ghisellini G., Villata M., De Francesco G., Lanteri L., Chiaberge M., Peila A., Antico G. Astron. Astrophys. Suppl. Ser. 127, 445 (1998) Optical photometric monitoring of gamma-ray loud blazars. II. Observations from November 1995 to June 1996

J/A+AS/127/463/  table1.dat table2.dat table3.dat table4.dat  (50Kb)

Fairall A.P., Woudt P.A., Kraan-Korteweg R.C. Astron. Astrophys. Suppl. Ser. 127, 463 (1998) Extragalactic Large-Scale Structures behind the Southern Milky Way. II. Redshifts Obtained at the SAAO in the Crux region

J/A+AS/127/471/  table2 table3  (67Kb)

Ortolani S., Bica E., Barbuy B. Astron. Astrophys. Suppl. Ser. 127, 471 (1998) V, I photometry of the bulge metal-rich globular clusters: NGC 6380 and Terzan 12

J/A+AS/127/521/  table1 table2 table3 table4 tables.ps tables.tex  (178Kb)

Wu H., Zou Z.L., Xia X.Y., Deng Z.G. Astron. Astrophys. Suppl. Ser. 127, 521 (1998) A statistical study of the spectra of very luminous IRAS galaxies. I. Data

J/A+AS/127/543/  table1  (81Kb)

Dimitrijevic M.S., Sahal-Brechot S. Astron. Astrophys. Suppl. Ser. 127, 543 (1998) Stark broadening of spectral lines of multicharged ions of astrophysical interest XVI: S V spectral lines

J/A+AS/128/19/  tab1-4 tab5-42 tab43-47  (7.8 Mb)

Kaluzny J., Kubiak M., Szymanski M., Udalski A., Krzeminski W., Mateo M., Stanek K.Z. Astron. Astrophys. Suppl. Ser. 128, 19 (1998) The Optical Gravitational lensing Experiment. Variable stars in globular clusters - IV. Fields 104A-E in 47 Tuc

J/A+AS/128/29/  table1 table2 table5  (52Kb)

Piters A.J.M., Van Paradijs J., Schmitt J.H.M.M. Astron. Astrophys. Suppl. Ser. 128, 29 (1998) X-ray/optical observations of stars with shallow convection zones (A8-G2 V)

J/A+AS/128/35/  table2 table3  (54Kb)

Sjouwerman L.O., van Langevelde H.J., Winnberg A., Habing H.J. Astron. Astrophys. Suppl. Ser. 128, 35 (1998) A new sample of OH/IR stars in the Galactic center

J/A+AS/128/67/  table1.dat table2.dat  (397Kb)

Slezak E., Durret F., Guibert J., Lobo C. Astron. Astrophys. Suppl. Ser. 128, 67 (1998) A photometric catalogue of galaxies in the cluster Abell 85

J/A+AS/128/131/  table1.dat table2.dat  (27Kb)

Claria J.J., Piatti A.E., Lapasset E. Astron. Astrophys. Suppl. Ser. 128, 131 (1998) Photometric study of the open cluster NGC 2323

J/A+AS/128/139/  table4 table4.tex  (35Kb)

Clement R., Fabregat J. Astron. Astrophys. Suppl. Ser. 128, 139 (1998) Four colour photometry of binary systems. III. The early-type binary system CR Cassiopeiae

J/A+AS/128/153/ notes.dat notes.ps notes.tex table3.dat table3a.ps table3a.tex table3b.ps table3b.tex (370Kb) Machalski J. Astron. Astrophys. Suppl. Ser. 128, 153 (1998) The revised GB/GB2 sample of extragalactic radio sources

J/A+AS/128/179/  table2a table2b  (88Kb)

Betlem H., ter Kuile C.R., de Lignie M., van’t Leven J., Jobse K., Miskotte K., Jenniskens P. Astron. Astrophys. Suppl. Ser. 128, 179 (1998) Precision meteor orbits obtained by the Dutch Meteor Society - Photographic Meteor Survey (1981-1993)

J/A+AS/128/203/  table1-3 table4 tables.tex  (62Kb)

Popovic L.C., Dimitrijevic M.S. Astron. Astrophys. Suppl. Ser. 128, 203 (1998) The electron impact broadening parameters in hot atmospheres: Mn II, Mn III, Ga III, Ge III and Ge IV lines

J/A+AS/128/245/  table2.dat table3.dat table4.dat table5.dat tablea.dat  (42Kb)

Adelman S.J. Astron. Astrophys. Suppl. Ser. 128, 245 (1998) uvby photometry of the chemically peculiar stars AR Aurigae, 3 Hydrae, 49 Cancri, and BX Boo

J/A+AS/128/251/  table2.dat table3.dat table4.dat  (98Kb)

Vaz L.P.R., Cunha N.C.S., Andersen J., Clausen J.V., Garcia J.M., Gimenez A., Casey B.W., de Koff S. Astron. Astrophys. Suppl. Ser. 128, 251 (1998) Four-colour photometry of eclipsing binaries. XXXVI. Light curves of the O7V+O9V system V3903 Sagittarii

J/A+AS/128/255/  table4.dat  (74Kb)

Su C.-G., Zhao J.-L., Tian K.-P. Astron. Astrophys. Suppl. Ser. 128, 255 (1998) Membership determination of stars using proper motions in the region of the open cluster M11

J/A+AS/128/265/  table2.dat table3.dat table7.dat  (22Kb)

Masana E., Jordi C., Maitzen H.M., Torra J. Astron. Astrophys. Suppl. Ser. 128, 265 (1998) CP2 stars as viewed by the ubvy-beta system

J/A+AS/128/299 See VII/206/

J/A+AS/128/309/  table1a.dat table1b.dat table2a.dat table2b.dat table4.dat table5.dat  (28Kb)

Kaltcheva N.T. Astron. Astrophys. Suppl. Ser. 128, 309 (1998) Stroemgren and Hβphotometry of OB stars in the region of the Carina Spiral Feature

J/A+AS/128/315/  table1.dat table1.tex  (38Kb)

Jia Guibin, Cen Xuefen, Ma Huiyu, Wang J. Astron. Astrophys. Suppl. Ser. 128, 315 (1998) Variability in Blazars

J/A+AS/128/331/ fig9a.ps fig9b.ps fig9c.ps fig9d.ps fig9e.ps fig9f.ps fig10a.ps fig10b.ps fig10c.ps fig10d.ps fig10e.ps fig10f.ps table2.dat table2.tex (5.1 Mb) Danner R. Astron. Astrophys. Suppl. Ser. 128, 331 (1998) Searching for Old Neutron Stars with ROSAT: I. Soft X-ray Sources in Molecular Clouds at High Galactic Latitude.

J/A+AS/128/359/  table1.dat  (160Kb)

Dimitrijevic M.S., Sahal-Brechot S. Astron. Astrophys. Suppl. Ser. 128, 359 (1998) Stark broadening of spectral lines of multicharged ions of astrophysical interest. XVII: Ca IX and Ca X spectral lines

J/A+AS/128/361/  rplines.dat rqplines.dat  (475Kb)

Roncin J.-Y., Launay F. Astron. Astrophys. Suppl. Ser. 128, 361 (1998) Table of vacuum ultraviolet emission band systems of molecular nitrogen from 82.6 to 124.2 nm

J/A+AS/128/409/  table5a.dat table5b.dat table6.dat  (3.6 Mb)

De Cuyper J.-P., Hensberge H. Astron. Astrophys. Suppl. Ser. 128, 409 (1998) Wavelength calibration at moderately high resolution

J/A+AS/128/417 See VII/199/

J/A+AS/128/457/  table2.dat  (37Kb)

Wroblewski H., Torres C. Astron. Astrophys. Suppl. Ser. 128, 457 (1998) New proper motion determination of Luyten catalogue stars (LTT) with declination between -5^{∘}and -30^{∘}and right ascension between 0h and 13h 30m.

J/A+AS/128/459/  table3.dat tables.dat  (56Kb)

Makarova L., Karanchentsev I., Takalo L.O., Heinamaki P., Valtonen, M. Astron. Astrophys. Suppl. Ser. 128, 457 (1998) Detailed images and distance measurements for eighteen irregular galaxies in the Canes Venatici cloud.

J/A+AS/128/471/ table1.dat table2.dat table3.dat table4.dat table5.dat table6.dat table7.dat table8.dat table9.dat table10.dat table11.dat table12.dat table13.dat table14.dat table15.dat table16.dat table17.dat table18.dat table19.dat (239Kb) Mowlavi N., Schaerer D., Meynet G., Bernasconi P.A., Charbonnel C. Maeder A. Astron. Astrophys. Suppl. Ser. 128, 471 (1998) Grids of stellar models. VII. From 0.8 to 60M_{⊙} at Z=0.10

J/A+AS/128/475/  table2.dat  (17Kb)

Jaschek C., Andrillat Y. Astron. Astrophys. Suppl. Ser. 128, 475 (1998) B[e] stars. V. HD 50138 = MWC 158

J/A+AS/128/485/ figure1a.ps figure1b.ps spectra/* table1.dat table2.dat table3.dat table4.dat table5.dat table6.dat (11.6 Mb) Montes D., Martin E.L. Astron. Astrophys. Suppl. Ser. 128, 485 (1998) Library of high-resolution UES echelle spectra of F, G, K and M field dwarf stars

J/A+AS/128/507/ fig4.ps fig5.ps fig6.ps fig7.ps fig8.ps fig9.ps fig10.ps fig11.ps fig12.ps fig13.ps fig14.ps fig15.ps table2-4.dat (990Kb) Engels D., Hagen H.-J., Cordis L., Koehler S., Wisotzki L., Reimers D. Astron. Astrophys. Suppl. Ser. 128, 507 (1998) + Erratum vol. 131, p. 196 The Hamburg Quasar Survey. II. A first list of 121 quasars

J/A+AS/128/545/  table3.dat table4.dat  (19Kb)

Rodrigues I., Dottori H., Cepa J., Vilchez J. Astron. Astrophys. Suppl. Ser. 128, 545 (1998) Physical properties of the ionized gas and brightness distribution in NGC 4736

J/A+AS/128/561/ fig1.ps fig2.ps fig3.ps fig4.ps fig5.ps fig6.ps fig7.ps fig8.ps fig9.ps fig10.ps fig11.ps table1.dat table2.dat (851Kb) Batalha C.C., Quast G.R., Torres C.A.O., Pereira P.C.R., Terra M.A.O., Jablonski F., Schiavon R.P., Sartori M.J. Astron. Astrophys. Suppl. Ser. 128, 561 (1998) Photometric variability of southern T Tauri stars

J/A+AS/128/573/  table3.dat table4.dat  (17Kb)

Maitzen H.M., Pressberger R., Paunzen E. Astron. Astrophys. Suppl. Ser. 128, 573 (1998) Δ-a and Stroemgren photometry of stars in the Renson-catalogue of Ap and Am stars

J/A+AS/129/1/  table1.dat  (15Kb)

Sazonov S.Y., Sunyaev R.A., Terekhov O.V., Lund N., Brandt S., Castro-Tirado A.J. Astron. Astrophys. Suppl. Ser. 129, 1 (1998) GRANAT/WATCH catalogue of cosmic gamma-ray bursts: December 1989 to September 1994

J/A+AS/129/9/ kw045.dat kw204v.dat kw204y.dat kw207.dat kw284.dat kw323.dat kw445.dat (36Kb) Pena J.H., Peniche R., Hobart M.A., Rolland A., Lopez de Coca P., Paparo M., Parrao L., de la Cruz C., Olivares J.I., Costa V., Ibanoglu C., Ertan A.Y., Tumer O., Evren S., Tunca Z. Astron. Astrophys. Suppl. Ser. 129, 9 (1998) Photoelectric photometry and period analysis of selected delta Scuti stars in Praesepe

J/A+AS/129/41/  table1.dat  (69Kb)

Allende Prieto C., Garcia Lopez R. J. Astron. Astrophys. Suppl. Ser. 129, 41 (1998) Fe I line shifts in the optical spectrum of the Sun

J/A+AS/129/83 See V/99/

J/A+AS/129/87/  notes.dat table3.dat  (79Kb)

Boller T., Bertoldi F., Dennefeld M., Voges W. Astron. Astrophys. Suppl. Ser. 129, 87 (1998) ROSAT All-Sky Survey observations of IRAS galaxies. I. Soft X-ray and far-infrared properties.

J/A+AS/129/147/  table3.dat  (21Kb)

Quinet P. Astron. Astrophys. Suppl. Ser. 129, 147 (1998) Radiative transition probabilities for forbidden lines in singly ionized cobalt (Co II)

J/A+AS/129/155/  table1.dat  (235Kb)

Dimitrijevic M.S., Sahal-Brechot S. Astron. Astrophys. Suppl. Ser. 129, 155 (1998) Stark broadening of spectral lines of multicharged ions of astrophysical interest XVIII: Si XI and Si XIII spectral lines

J/A+AS/129/237/  stars.dat table11.dat  (46Kb)

Feltzing S., Gustafsson B. Astron. Astrophys. Suppl. Ser. 129, 237 (1998) Abundances of metal-rich stars. Detailed abundance analysis of 47 G and K dwarf stars with [Me/H] 0.10 dex.

J/A+AS/129/281/  catalog.dat  (54Kb)

Durret F., Felenbok P., Lobo C., Slezak E. Astron. Astrophys. Suppl. Ser. 129, 281 (1998) A catalogue of velocities in the cluster of galaxies Abell 85

J/A+AS/129/289/  fig/* refs.dat stars.dat table1.dat  (2.0 Mb)

Moujtahid A., Zorec J. Astron. Astrophys. Suppl. Ser. 129, 289 (1998) Long-term visual spectrophotometric behaviour of Be stars

J/A+AS/129/337/  tables.dat  (19Kb)

Harder S., Bertout C., Mennessier C. Astron. Astrophys. Suppl. Ser. 129, 337 (1998) CCD Photometry of YY Orionis stars

J/A+AS/129/357/  appena.dat table2.dat  (74Kb)

Faundez-Abans M., de Oliveira-Abans M. Astron. Astrophys. Suppl. Ser. 129, 357 (1998) On the morphology of peculiar ring galaxies.

J/A+AS/129/399/  enacs.dat table2.dat table5.dat  (433Kb)

Katgert P., Mazure A., den Hartog R., Adami C., Biviano A., Perea J. Astron. Astrophys. Suppl. Ser. 129, 399 (1998) The ESO Nearby Abell Cluster Survey. V. The catalogue: contents and instructions for use

J/A+AS/129/431 See II/215/

J/A+AS/129/455/ t4n520.dat t5n772.dat t6n1023.dat t7n1253.dat t8a141.dat t9n2535.dat t10n2623.dat t11n2782.dat t12n3226.dat t13n3656.dat t14n3690.dat t15n3921.dat (124Kb)

Deeg H.J., Munoz-Tunon C., Tenorio-Tagle G., Telles E., Vilchez J.M., Rodriguez-Espinosa J.M., Duc P.A., Mirabel I.F Astron. Astrophys. Suppl. Ser. 129, 455 (1998) A catalogue of dwarf galaxy candidates around interacting galaxies

J/A+AS/129/505/  table1.dat  (15Kb)

Blackwell D.E., Lynas-Gray A.E. Astron. Astrophys. Suppl. Ser. 129, 505 (1998) Determination of the temperatures of selected ISO flux calibration stars using the Infrared Flux Method

J/A+AS/129/541/  tableb1.dat  (44Kb)

Oudmaijer R.D. Astron. Astrophys. Suppl. Ser. 129, 541 (1998) High resolution spectroscopy of the post-red supergiant IRC +10420. I: The data.

J/A+AS/129/563/  abund.dat table3.dat table3.ps table3.tex  (151Kb)

Pintado O.I., Adelman S.J., Gulliver A.F. Astron. Astrophys. Suppl. Ser. 129, 563 (1998) Elemental abundance analyses with Complejo Astronomico El Leoncito REOSC Echelle Sprectrograms. III. HR 4487, 14 Hydrae and 3 Centauri A

J/A+AS/129/569/  table5.dat  (10Kb)

Jia Guibin, Cen Xuefen, Ma Huiyu, Wang Jiancheng Astron. Astrophys. Suppl. Ser. 129, 569 (1998) Microvariability of the blazar OJ 287

J/A+AS/129/577/  table2.dat table2.tex  (14Kb)

Takalo L.O., Sillanpaa A., Valtaoja E., Katajainen S., Nilsson K. Pursimo T., Heinamaki P., Villata M., Raiteri C.M., Ghisellini G., De Francesco G., Lanteri L., Chiaberge M., Peila A., Boltwood P., Tosti G., Fiorucci M., Terasranta H., Tornikoski M., Teerikorpi P. Astron. Astrophys. Suppl. Ser. 129, 577 (1998) Monitoring of AO 0235+164 during a faint state

J/A+AS/129/583/  name.dat table1.dat  (382Kb)

Fasano G., Filippi M., Bertola F. Astron. Astrophys. Suppl. Ser. 129, 583 (1998) Surface photometry of early-type galaxies in the Hubble Deep Field

J/A+AS/129/593/  fig1.dat  (207Kb)

Cunow B. Astron. Astrophys. Suppl. Ser. 129, 593 (1998) BVRI surface photometry for a sample of 14 Sb galaxies

J/A+AS/129/603/  table2.dat table3.dat table4.dat table5.dat table6.dat table7.dat table8.dat  (38Kb)

Quinet P. Astron. Astrophys. Suppl. Ser. 129, 603 (1998) Predicted term energies, wavelengths and oscillator strengths for transitions involving high-nl Rydberg states in C II, C III and C IV

J/A+AS/130/65/  ( total of 7 files)  (8.8 Mb)

Lejeune T., Cuisinier F., Buser R. Astron. Astrophys. Suppl. Ser. 130, 65 (1998) A standard stellar library for evolutionary synthesis: II. The M dwarf extension

J/A+AS/130/77/  table7.dat  (15Kb)

Jones D.H.P., Taylor D.B., Williams I.P. Astron. Astrophys. Suppl. Ser. 130, 77 (1998) Astrometry of Satellites I to V of Uranus

J/A+AS/130/109/  table2.dat  (31Kb)

Tosti G., Fiorucci M., Luciani M., Rizzi M. Villata N., Raiteri C.M., De Francesco G., Lanteri L., Chiaberge M., Peila A., Cavallone M., Sobrito G., Maesano M., Massaro E., Montagni F., Nesci R., Ghisellini G., Takalo L.O., Sillanpaa A., Katajainen S., Heinamaki P., Nilsson K., Pursimo T. Astron. Astrophys. Suppl. Ser. 130, 109 (1998) BV(RI)c monitoring of ON 231 during the great outburst in 1994-1997

J/A+AS/130/117/  table1.dat table2.dat  (16Kb)

Docobo J.A. Astron. Astrophys. Suppl. Ser. 130, 117 (1998) Micrometer measurements of double stars from the Spanish observatories at Calar Alto and Santiago de Compostela

J/A+AS/130/133/  table1.dat  (98Kb)

Reiz A., Franco G.A.P. Astron. Astrophys. Suppl. Ser. 130, 133 (1998) UBV polarimetry of 361 A- and F-type stars in Selected Areas

J/A+AS/130/157/  table2.dat  (105Kb)

van Leeuwen F., Evans D.W. Astron. Astrophys. Suppl. Ser. 130, 157 (1998) On the use of the Hipparcos intermediate astrometric data

J/A+AS/130/223/ table6.dat table7.dat table8.dat table9.dat table10.dat table11.dat table12.dat (25Kb) North P., Carquillat J.-M., Ginestet N., Carrier F., Udry S. Astron. Astrophys. Suppl. Ser. 130, 223 (1998) Multiplicity among peculiar A stars I. The Ap stars HD 8441 and HD 137909, and the Am stars HD 43478 and HD 96391

J/A+AS/130/233/  catalog.dat table1.dat  (133Kb)

Crosby N., Lund N., Vilmer N., Sunyaev R. Astron. Astrophys. Suppl. Ser. 130, 233 (1998) The WATCH Solar X-Ray Burst Catalogue

J/A+AS/130/245/  table2.dat  (118Kb)

Vaz L.P.R., Andersen J., Casey B.W., Clausen J.V., Mathieu R.D., Heyer I. Astron. Astrophys. Suppl. Ser. 128, 251 (1998) Four-colour photometry of eclipsing binaries. XXXIX. Light curves of the pre-main sequence triple system TY Coronae Australis

J/A+AS/130/251/ table1.dat table2.dat table5.dat table6.dat table8.dat table9.dat table10.dat table11.dat table12.dat table13.dat table14.dat table15.dat (86Kb) Longhetti M., Rampazzo R., Bressan A., Chiosi C. Astron. Astrophys. Suppl. Ser. 130, 251 (1998) Star formation history of early-type galaxies in low density environments. I. Nuclear line-strength indices.

J/A+AS/130/285/  table1.dat table3.dat table4.dat table5.dat table6.dat  (86Kb)

Contini T., Considere S., Davoust E. Astron. Astrophys. Suppl. Ser. 130, 285 (1998) Starbursts in barred spiral galaxies. III. Definition of a homogeneous sample of Starburst Nucleus Galaxies

J/A+AS/130/323/  table2.dat table3.dat  (325Kb)

Vettolani G., Zucca E., Merighi R., Mignoli M., Proust D., Zamorani G., Cappi A., Guzzo L., Maccagni D., Ramella M., Stirpe G.M., Blanchard A., Cayatte V., Collins C., McGillivray H., Maurogordato S., Scaramella R., Balkowski C., Chincarini G., Felenbok P. Astron. Astrophys. Suppl. Ser. 130, 323 (1998) The ESO Slice Project (ESP) galaxy redshift survey: III. The Sample

J/A+AS/130/333/  appena.dat appena.ps appena.tex fig5/* table2.dat  (1.9 Mb)

Theureau G., Bottinelli L., Coudreau-Durand N., Gouguenheim L., Hallet N., Loulergue M., Paturel G., Teerikorpi P. Astron. Astrophys. Suppl. Ser. 130, 333 (1998) Kinematics of the local universe. VII. New 21-cm line measurements of 2112 galaxies

J/A+AS/130/341/  table1.dat  (20Kb)

Trasarti-Battistoni R. Astron. Astrophys. Suppl. Ser. 130, 341 (1998) Loose Groups of Galaxies in the Perseus-Pisces Survey

J/A+AS/130/359/  table3.dat table4.dat tables.tex  (45Kb)

Spagna A., Lattanzi M.G., McLean B.J., Massone G., Lasker B.M. Astron. Astrophys. Suppl. Ser. 130, 359 (1998) Faint photometric BVRc CCD sequences. The North Galactic Pole (b=85 deg) and the anticenter (l=133.2, b=-1.6 deg)

J/A+AS/130/421/  refs.dat table5a.dat table5b.dat table6a.dat table6b.dat  (155Kb)

Filipovic M.D., Haynes R.F., White G.L., Jones P.A. Astron. Astrophys. Suppl. Ser. 130, 421 (1998) A radio continuum study of the Magellanic Clouds. Part VII: Discrete radio sources in the Magellanic Clouds

J/A+AS/130/455/  table2.dat table2.ps table2.tex  (272Kb)

Vogt N., Kerschbaum F., Maitzen H.M., Faundez-Abans M. Astron. Astrophys. Suppl. Ser. 130, 455 (1998) An extensive Δa-photometric survey of southern B and A type bright stars

J/A+AS/130/477/  table1.dat table2.dat table3.dat tables.ps tables.tex  (137Kb)

Knude J. Astron. Astrophys. Suppl. Ser. 130, 477 (1998) Stellar uvbyβphotometry in three EUV shadow directions

J/A+AS/130/513/ table7.dat table8.dat table9.dat table10.dat table12.dat table34.dat table56.dat tables.ps tables.tex (161Kb) Garcia-Vargas M.L., Molla M., Bressan A. Astron. Astrophys. Suppl. Ser. 130, 513 (1998) Calcium triplet synthesis

J/A+AS/130/539/  table1.dat  (185Kb)

Dimitrijevic M.S., Sahal-Brechot S. Astron. Astrophys. Suppl. Ser. 130, 539 (1998) Stark broadening of spectral lines of multicharged ions of astrophysical interest. XIX: Na X spectral lines

J/A+AS/130/541/  table2.dat table7.dat  (1.0 Mb)

Raassen A.J.J., Pickering J.C., Uylings P.H.M. Astron. Astrophys. Suppl. Ser. 130, 541 (1998) Calculated oscillator strengths of singly ionized Cobalt

J/A+AS/131/1/  tables.dat  (54Kb)

Karachentsev I.D., Drozdovsky I.O. Astron. Astrophys. Suppl. Ser. 131, 1 (1998) Distances to five resolved galaxies in the CVn cloud

J/A+AS/131/11/  name.dat table2.dat  (126Kb)

Reich W., Reich P., Pohl M., Kothes R., Schlickeiser R. Astron. Astrophys. Suppl. Ser. 131, 11 (1998) Flux density monitoring of radio sources with detected or supposed gamma-ray emission

J/A+AS/131/73/  hidata.dat notes.dat opdata.dat refs.dat  (442Kb)

Martin M.C. Astron. Astrophys. Suppl. Ser. 131, 73 (1998) Catalogue of HI maps of galaxies. I.

J/A+AS/131/77/  refs.dat table1a.dat table1b.dat  (61Kb)

Martin M.C. Astron. Astrophys. Suppl. Ser. 131, 77 (1998) Catalogue of HI maps of galaxies. II. Analysis of the data

J/A+AS/131/89/  table5.dat  (47Kb)

Tian K.-P., Zhao J.-L., Shao Zh.-Y., Stetson P.B. Astron. Astrophys. Suppl. Ser. 131, 89 (1998) Determination of proper motions and membership of the open clusters NGC 1750 and NGC 1758.

J/A+AS/131/105/  table3.dat table4.dat table5.dat table6.dat table7.dat table8.dat  (45Kb)

Vega Beltran J.C., Zeilinger W.W., Amico P., Schultheis M., Corsini E.M., Funes J.G., Beckman J., Bertola F. Astron. Astrophys. Suppl. Ser. 131, 105 (1998) Mixed early and late-type properties in the bar of NGC 6221: evidence for evolution along the Hubble sequence?

J/A+AS/131/115/  table2a.dat table2b.dat table345.dat tables.tex  (110Kb)

Li Z.-Y., Ding Y.-R., Zhang Z.-S. Li Y.-I. Astron. Astrophys. Suppl. Ser. 131, 115 (1998) The variations in light and color curves and possible short period oscillations of binary AU Serpentis

J/A+AS/131/119/  table1.dat  (20Kb)

Cieslinski D., Steiner J.E., Jablonski F.J. Astron. Astrophys. Suppl. Ser. 131, 119 (1998) Southern and equatorial irregular variables. II. Optical spectroscopy

J/A+AS/131/137/  table2.dat table3.dat table4.dat table5.dat  (19Kb)

Chandra S., Rashmi Astron. Astrophys. Suppl. Ser. 131, 137 (1998) Einstein A-Coefficients for rotational transitions in the ground vibrational state of ^{28}SiC_{2}, ^{29}SiC_{2} and ^{30}SiC_{2}

J/A+AS/131/141/  table1.dat table2.dat  (105Kb)

Dimitrijevic M.S., Sahal-Brechot S. Astron. Astrophys. Suppl. Ser. 131, 141 (1998) Stark broadening of spectral lines of multicharged ions of astrophysical interest. XX: O VII and Mg XI spectral

lines

J/A+AS/131/143/  table1.dat table2.dat table3.dat table4.dat  (152Kb)

Dimitrijevic M.S., Sahal-Brechot S. Astron. Astrophys. Suppl. Ser. 131, 143 (1998) Stark broadening of spectral lines of multicharged ions of astrophysical interest. XXI: Sc X, Sc XI, Ti XI and Ti XII spectral lines

J/A+AS/131/153/ table1.dat table3.dat table4.dat table5.dat table6.dat table7.dat table8.dat table9.dat tables.ps tables.tex (224Kb) Binello A.M., Mason H.E., Storey P.J. Astron. Astrophys. Suppl. Ser. 131, 153 (1998) Atomic data from the IRON Project. XXXI. Electron impact excitation of optically allowed and intercombination electric dipole transitions in FeXII

J/A+AS/131/197/  notes.dat table3.dat  (35Kb)

Metanomski A.D.F., Pasquini L., Krautter J., Cutispoto G., Fleming T.A. Astron. Astrophys. Suppl. Ser. 131, 197 (1998) F, G and K stars in the ROSAT all-sky survey. I. Photometry.

J/A+AS/131/209/  table1.dat table2.dat table3.dat  (35Kb)

Alonso A., Arribas S., Martinez-Roger C. Astron. Astrophys. Suppl. Ser. 131, 209 (1998) Broad-band JHK(L’) photometry of a sample of giants with 0.5 [Fe/H] -3

J/A+AS/131/221/  table1.dat  (13Kb)

Solanes J.M., Stein P. Astron. Astrophys. Suppl. Ser. 131, 221 (1998) Kinematics of the southern galaxy cluster Abell 3733

J/A+AS/131/239/  notes.dat table6.dat table11.dat  (10.3 Mb)

Galadi-Enriquez D., Jordi C., Trullols E., Guibert J., Tian K.P., Zhao J.L. Astron. Astrophys. Suppl. Ser. 131, 239 (1998) The overlapping open clusters NGC 1750 and NGC 1758: II. BVR photographic photometry and proper motions

J/A+AS/131/287/  table1.dat table2.dat table3.dat  (21Kb)

Simien F., Prugniel P. Astron. Astrophys. Suppl. Ser. 131, 287 (1998) Kinematical data on early-type galaxies. IV.

J/A+AS/131/291/  table1.dat  (32Kb)

Veiga C.H., Vieira Martins R. Astron. Astrophys. Suppl. Ser. 131, 291 (1998) New CCD positions of Triton and a comparison with the theoretical predictions

J/A+AS/131/295/  table2.dat table3.dat table4.dat tables.ps tables.tex  (124Kb)

Ugryumov A.V., Pustilnik S.A., Lipovetsky V.A., Izotov Yu.I., Richter G. Astron. Astrophys. Suppl. Ser. 131, 295 (1998) Spectral survey of Case emission-line galaxies with the 6m Russian telescope

J/A+AS/131/303/  table4.dat table5.dat tables.tex  (54Kb)

Stanghellini C., O’Dea C.P., Dallacasa D., Baum S.A., Fanti R., Fanti C. Astron. Astrophys. Suppl. Ser. 131, 303 (1998) A complete sample of GHz-Peaked-Spectrum radio sources and its radio properties

J/A+AS/131/317/  catalog.dat  (26Kb)

Petit H. Astron. Astrophys. Suppl. Ser. 131, 317 (1998) Catalogue of HII regions in NGC 7331 at the 6 meter telescope

J/A+AS/131/319/  catalog.dat  (185Kb)

Petit H., Figon P., Petit M. Astron. Astrophys. Suppl. Ser. 131, 319 (1998) Catalogue of HII regions measured on 6m telescope plates at Observatoire de Marseille

J/A+AS/131/321/  table1.dat table3.dat table4.dat tables.ps tables.tex ubvri.dat  (147Kb)

Cutispoto G. Astron. Astrophys. Suppl. Ser. 131, 321 (1998) Long-term monitoring of active stars. VIII. UBV(RI)c photometry collected in February 1992

J/A+AS/131/395/ init.dat ld1.dat ld2a.dat ld2b.dat ld2c.dat ld2d.dat models.dat table01.txt table02.txt table03.txt table04.txt table05.txt table06.txt table07.txt table08.txt table09.txt table10.txt table11.txt table12.txt table13.txt table14.txt table15.txt table16.txt table17.txt table18.txt table19.txt table20.txt table21.txt table22.txt table23.txt table24.txt (953Kb) Claret A. Astron. Astrophys. Suppl. Ser. 131, 395 (1998) Comprehensive tables for the interpretation and modeling of the light curves of eclipsing binaries

J/A+AS/131/401/ position.dat tab6-11.ps tab6-11.tex tab12-16.ps tab12-16.tex tab17-22.ps tab17-22.tex tables.dat (366Kb) Yudin R.V., Evans A. Astron. Astrophys. Suppl. Ser. 131, 401 (1998) Polarimetry of southern peculiar early-type stars

J/A+AS/131/431/  table1.dat  (240Kb)

Allende Prieto C., Garcia Lopez R.J. Astron. Astrophys. Suppl. Ser. 131, 431 (1998) A catalogue of accurate wavelengths in the optical spectrum of the Sun

J/A+AS/131/469/  table1.dat table2.dat table3.dat tables.ps tables.tex  (154Kb)

Miroshnichenko A.S., Fremat Y., Houziaux L., Andrillat Y., Chentsov E.L., Klochkova V.G. Astron. Astrophys. Suppl. Ser. 131, 469 (1998) High resolution spectroscopy of the galactic candidate LBV MWC 314

J/A+AS/131/479/  table2.dat  (13Kb)

Andrillat Y., Jaschek C. Astron. Astrophys. Suppl. Ser. 131, 479 (1998) B[e] stars. VI. MWC 297 = IRAS 18250-0351

J/A+AS/131/499/  acds.dat  (79Kb)

Galavis M.E., Mendoza C., Zeippen C.J. Astron. Astrophys. Suppl. Ser. 131, 499 (1998) Atomic data from the IRON Project. XXIX. Radiative rates for transitions within the n=2 complex in ions of the boron isoelectronic sequence

J/A+AS/132/13/  table2.dat table2.ps  (447Kb)

Durand S., Acker A., Zijlstra A. Astron. Astrophys. Suppl. Ser. 132, 13 (1998) The kinematics of 867 galactic planetary nebulae

J/A+AS/132/29/  table1.dat  (8Kb)

Thorel J.-C. Astron. Astrophys. Suppl. Ser. 132, 29 (1998) Mesures d’etoiles doubles effectuees a la lunette de 50 cm de l’Observatoire de Nice (2^{e} serie). Measures of visual double stars made with the 50 cm refractor at the Nice Observatory (2^{nd} series).

J/A+AS/132/31/  refs.dat table1.dat table2.dat  (162Kb)

Nilsson K. Astron. Astrophys. Suppl. Ser. 132, 31 (1998) Kinematical models for double radio sources and the unified scheme. II. The database

J/A+AS/132/83/  table2.dat  (20Kb)

Bai J.M, Xie G.Z., Li K.H., Zhang X., Liu W.W. Astron. Astrophys. Suppl. Ser. 132, 83 (1998) The intraday variability in the radio-selected and X-ray-selected BL Lac objects

J/A+AS/132/93/  table1.dat table2.dat table3.dat  (28Kb)

Adelman S.J. Astron. Astrophys. Suppl. Ser. 132, 93 (1998) On the Hipparcos photometry of Chemically Peculiar B, A, and F stars

J/A+AS/132/155/  table2.dat table2.ps table2.tex  (423Kb)

Huensch M., Schmitt H.H.M.M., Voges W. Astron. Astrophys. Suppl. Ser. 132, 155 (1998) The ROSAT all-sky survey catalogue of optically bright main-sequence stars and subgiant stars

J/A+AS/132/173/  table2.dat table3.dat table4.dat table5.dat tables.ps tables.tex  (75Kb)

Li J.Z., Hu J.Y. Astron. Astrophys. Suppl. Ser. 132, 173 (1998) Newly discovered candidate weak-line T Tauri stars in the surrounding area of the Taurus-Auriga region

J/A+AS/132/181/  table1.dat table3.dat tables.ps tables.tex  (145Kb)

Wu H., Zou Z.L., Xia X.Y., Deng Z.G. Astron. Astrophys. Suppl. Ser. 132, 181 (1998) A statistical study of the spectra of very luminous IRAS galaxies. II. Spectral and environmental analysis

J/A+AS/132/195/  table2.dat table3.dat  (84Kb)

Noel F. Astron. Astrophys. Suppl. Ser. 132, 195 (1998) Observations of the Sun with the astrolabe of Santiago: 1995-1997

J/A+AS/132/211/ fig/* onsa2-1.dat onsacat.dat sest2-1.dat sest3-2.dat sestcat.dat tableb1.ps tableb1.tex tableb2.ps tableb2.tex (1.3 Mb) Harju J., Lehtinen K., Booth R.S., Zinchenko I. Astron. Astrophys. Suppl. Ser. 132, 211 (1998) A survey of SiO emission towards interstellar masers. I. SiO line characteristics.

J/A+AS/132/233/  appendix.dat  (21Kb)

Xu J., Li D., Liu J., Cheng H., Zhang Z., Raschotin S.I., Emeliyanov V.A., Kaplenko V.V., Zalutsky V.T. Astron. Astrophys. Suppl. Ser. 132, 233 (1998) Continuous declination system set up by observations of photoelectric astrolabe Mark I in Irkutsk. The first results of international cooperation between CSAO and VS NIIFTRI.

J/A+AS/132/237/  table1.dat table2.dat table3.dat table4.dat  (75Kb)

Alzner A. Astron. Astrophys. Suppl. Ser. 132, 237 (1998) Measurements of double stars 1993.67-1998.13

J/A+AS/132/253/ bu151s.dat bu151v.dat hu39h.dat hu39l.dat stf1130.dat stf2434.dat stt53.dat stt89.dat stt229.dat stt363.dat stt448.dat table.dat (53Kb) Alzner A. Astron. Astrophys. Suppl. Ser. 132, 253 (1998) Orbital elements of 9 visual double stars

J/A+AS/132/255/  table1.dat table2.dat table3.dat table4.dat  (348Kb)

Golev V., Prugniel P. Astron. Astrophys. Suppl. Ser. 132, 255 (1998) A catalogue of Mg_{2} indices of galaxies and globular clusters

J/A+AS/132/305/  table3.dat table4.dat  (691Kb)

Terasranta H., Tornikoski M., Mujunen A., Karlamaa K., Valtonen T., Henelius N., Urpo S., Lainela M., Pursimo T., Nilsson K., Wiren S., Lahteenmaki A., Korpi M., Rekola R., Heinamaki P., Hanski M., Nurmi P., Kokkonen K., Keinanen P., Joutsamo O., Oksanen J., Pietila H., Valtaoja E., Valtonen M., Kononen P. Astron. Astrophys. Suppl. Ser. 132, 305 (1998) Fifteen years monitoring of extragalactic radio sources at 22, 37 and 87GHz

J/A+AS/132/341/  table4.dat table5.dat  (25Kb)

Motch C., Guillout P., Haberl F., Krautter J., Pakull M.W., Pietsh W., Reinsch K., Voges W., Zickgraf F.-J. Astron. Astrophys. Suppl. Ser. 132, 341 (1998) Identification of selected sources for the ROSAT Galactic Plane Survey. I.

J/A+AS/132/361/  mean.dat tab2-9.dat  (12Kb)

Raiteri C.M., Villata M., De Francesco G., Lanteri L., Cavallone M., Sobrito G. Astron. Astrophys. Suppl. Ser. 132, 361 (1998) Optical observations of eight X-ray selected BL Lacertae objects

J/A+AS/132/367/  table1.dat  (36Kb)

Barblan F., Bartholdi P., North P., Burki G., Olson E.C. Astron. Astrophys. Suppl. Ser. 132, 367 (1998) Physical parameters of the Algol system TZ Eridani from simultaneous analysis of Geneva 7-colour light curves

J/A+AS/133/1/  table1.dat  (17Kb)

Paunzen E., Maitzen H.M. Astron. Astrophys. Suppl. Ser. 133, 1 (1998) New variable chemically peculiar stars identified in the Hipparcos archive

J/A+AS/133/7/  table1.dat table1.tex  (22Kb)

Niarchos P.G, Mantegazza L., Poretti E., Manimanis V. Astron. Astrophys. Suppl. Ser. 133, 7 (1998) YY CMi: contact or near contact system?

J/A+AS/133/25/  table1.dat table2.dat table3.dat  (398Kb)

Kaluzny J.

Astron. Astrophys. Suppl. Ser. 133, 25 (1998) CCD photometry of distant open clusters. IV. Trumpler 5.

J/A+AS/133/29/  tab10-17.dat table18.dat tables.ps tables.tex  (321Kb)

Hatchell J., Thompson M.A., Millar T.J., Macdonald G.H. Astron. Astrophys. Suppl. Ser. 133, 29 (1998) A survey of molecular line emission towards ultracompact HII regions

J/A+AS/133/123 See VI/118/

J/A+AS/133/129/  table1.dat  (56Kb)

Gregorini L., Vigotti M., Mack K.-H., Zonnchen J., Klein U. Astron. Astrophys. Suppl. Ser. 133, 129 (1998) Multi-Frequency Study of the B3-VLA Sample. I. 10.6 GHz Data.

J/A+AS/133/171/  charts.ps table.dat table.ps table.tex  (8.7 Mb)

Surace C., Comte G. Astron. Astrophys. Suppl. Ser. 133, 171 (1998) The Marseille Schmidt survey for active star-forming galaxies. I. Data on 92 emission line objects in two fields

J/A+AS/133/181/  list.dat tab2-8.dat tab9-17.dat  (41Kb)

Makarova L., Karachentsev I.D. Astron. Astrophys. Suppl. Ser. 133, 181 (1998) CCD imaging of twenty nearby isolated irregular galaxies

J/A+AS/133/197/  table2.dat table3.dat table4.dat table5.dat tablea.dat  (51Kb)

Adelman S.J., Pi C.-L.M., Rayle K. Astron. Astrophys. Suppl. Ser. 133, 197 (1998) uvby photometry of 33 Tauri, HD 50169, and HR 7786 and an assessment of FCAPT comparison stars

J/A+AS/133/201/  table6.dat table6.ps table6.tex  (69Kb)

Iyengar K.V.K., MacConnell D.J. Astron. Astrophys. Suppl. Ser. 133, 201 (1998) Spectral classification of unidentified IRAS point sources with F_{ν}(12μm) = F_{ν}(25μm).

J/A+AS/133/221/  spectra/* stars.dat table1.dat  (73.7 Mb)

Soubiran C., Katz D., Cayrel R. Astron. Astrophys. Suppl. Ser. 133, 221 (1998) On line determination of stellar atmospheric parameters Teff, log g, [Fe/H] from ELODIE echelle spectra. II - The library of F5 to K7 stars.

J/A+AS/133/227/  table1.dat table2.dat  (108Kb)

Dimitrijevic M.S., Sahal-Brechot S. Astron. Astrophys. Suppl. Ser. 133, 227 (1998) Stark broadening of spectral lines of multicharged ions of astrophysical interest. XXII: K VIII and K IX spectral lines

J/A+AS/133/257/  table1.dat table2.dat table3.dat table4.dat table5.dat table6.dat  (83Kb)

Kisielius R., Storey P.J., Davey A.R., Neale L.T. Astron. Astrophys. Suppl. Ser. 133, 257 (1998) Recombination coefficients for Ne II lines at nebular temperatures and densities

J/A+AS/133/299/  table1.dat table1.ps table1.tex  (326Kb)

Damineli A., Stahl O., Kaufer A., Wolf B., Quast G., Lopes D.F. Astron. Astrophys. Suppl. Ser. 133, 299 (1998) Long-term spectroscopy of eta Carinae. I. The high and low excitation phases

J/A+AS/133/317/  table1.dat table2.dat table3.dat  (46Kb)

Heraudeau P., Simien F. Astron. Astrophys. Suppl. Ser. 133, 317 (1998) Stellar kinematical data for the central region of spiral galaxies

J/A+AS/133/325/  table3.dat  (40Kb)

Graham A.W., Colless M.M., Busarello G., Zaggia S., Longo G. Astron. Astrophys. Suppl. Ser. 133, 325 (1998) Extended stellar kinematics of elliptical galaxies in the Fornax cluster

J/A+AS/133/337/  table1.dat table2.dat  (19Kb)

Zinchenko I., Pirogov L., Toriseva M. Astron. Astrophys. Suppl. Ser. 133, 337 (1998) Studies of dense molecular cores in regions of massive star formation. VII. Core properties on the galactic scale

J/A+AS/133/387/  table5.dat  (72Kb)

Balaguer-Nunnez L., Tian K.P., Zhao J.L. Astron. Astrophys. Suppl. Ser. 133, 387 (1998) Determination of proper motions and membership of the open clusters NGC 1817 and NGC 1807

J/A+AS/133/403/  table1.dat table2.dat tables.ps tables.tex  (521Kb)

Mazzotta P., Mazzitelli G., Colafrancesco S., Vittorio N. Astron. Astrophys. Suppl. Ser. 133, 403 (1998) Ionization balance for optically thin plasmas: rate coefficients for all atoms and ions of the elements H to Ni.

J/A+AS/134/53/ table3.dat table4a.dat table4b.dat table5a.dat table5d.dat table6.dat tablea.dat (44Kb) Adelman S.J. Astron. Astrophys. Suppl. Ser. 134, 53 (1999) uvby Photometry of the chemically peculiar stars. HD 15980, HR 1094, 33 Gem, and HD 115708

J/A+AS/134/59/  appendxa.dat appendxb.dat  (114Kb)

Hilker M., Kissler-Patig M., Richtler T., Infante L., Quintana H. Astron. Astrophys. Suppl. Ser. 134, 59 (1999) The central region of the Fornax cluster. I. A catalog and photometric properties of galaxies in selected CCD fields

J/A+AS/134/75/  table2.dat  (15Kb)

Hilker M., Infante L., Vieira G., Infante L., Kissler-Patig M., Richtler T. Astron. Astrophys. Suppl. Ser. 134, 75 (1999) The central region of the Fornax cluster. II. Spectroscopy and radial velocities of member and background galaxies

J/A+AS/134/87/  table1.dat  (15Kb)

Jasinta D.M.D., Raharto M., Soegiartini E. Astron. Astrophys. Suppl. Ser. 134, 87 (1999) Photographic observations of visual double stars

J/A+AS/134/115/  table2.dat  (21Kb)

Slysh V.I., Val’tts I.E., Kalenskii S.V., Voronkov M.A., Palagi F., Tofani G., Catarzi M. Astron. Astrophys. Suppl. Ser. 134, 115 (1999) The Medicina survey of methanol masers at 6.7 GHz

J/A+AS/134/221/  notes.dat table1.dat table2.dat table3.dat tables.tex  (69Kb)

Brinkmann W., Chester M., Kollgaard R., Feigelson E., Voges W., Hertz P. Astron. Astrophys. Suppl. Ser. 134, 221 (1999) Radio-loud ROSAT sources near the North Ecliptic Pole.

J/A+AS/134/241/  notes.txt refs.dat table1.dat  (48Kb)

Phillips J.P. Astron. Astrophys. Suppl. Ser. 134, 241 (1999) Rotation in molecular clouds

J/A+AS/134/287/  table3.dat table4.dat table5.dat tables.ps tables.tex  (627Kb)

Reynolds A.P., Parmar A.N., Hakala P.J., Pollock A.M.T., Williams O.R., Peacock A., Taylor B.G. Astron. Astrophys. Suppl. Ser. 134, 287 (1999) The EXOSAT Medium energy Slew Survey Catalog

J/A+AS/134/301/  table7.dat  (7Kb)

Claria J.J., Mermilliod J.-C., Piatti A.E. Astron. Astrophys. Suppl. Ser. 134, 301 (1999) Membership, binarity and metallicity of red giants in the southern open cluster NGC 2354

J/A+AS/134/401/  table1.dat table2.dat  (28Kb)

Jackson N., Roland J., Bremer M., Rhee G., Webb J. Astron. Astrophys. Suppl. Ser. 134, 401 (1999) VLA observations of a 4C equatorial sample. II.

J/A+AS/134/463/  tablea1.dat tablea1.ps tablea1.tex  (29Kb)

Florczak M., Lazzaro D., Mothe-Diniz T., Angeli C.A., Betzler A.S. Astron. Astrophys. Suppl. Ser. 134, 463 (1999) A Spectroscopic Study of the Themis Family

J/A+AS/134/483/  table3.dat table3.ps table3.tex  (73Kb)

Hagen H.-J., Engels D., Reimers D. Astron. Astrophys. Suppl. Ser. 134, 483 (1999) The Hamburg Quasar Survey. III. Further new bright quasars.

J/A+AS/134/489/ ngc330.dat ngc346.dat ngc1818.dat ngc1948.dat ngc2004.dat ngc2100.dat tables.ps tables.tex (106Kb) Keller S.C., Wood P.R., Bessell M.S. Astron. Astrophys. Suppl. Ser. 134, 489 (1999) Be stars in and around young clusters in the Magellanic Clouds

J/A+AS/134/523/  catalog.dat comment.dat program.f refs.dat temper.dat unused.dat  (118Kb)

Taylor B.J. Astron. Astrophys. Suppl. Ser. 134, 523 (1999) Catalogs of temperatures and [Fe/H] averages for evolved G and K stars

J/A+AS/134/525/  catalog.dat  (285Kb)

Belikov A.N., Kharchenko N.V., Piskunov A.E., Schilbach E. Astron. Astrophys. Suppl. Ser. 134, 525 (1999) The extremely young open cluster NGC 6611. I. Compiled catalogue, absorption map and the HR diagram

J/A+AS/134/545/  refs.dat table1.dat  (18Kb)

Aristidi E., Prieur J.-L., Scardia M., Koechlin L., Avila R., Carbillet M., Lopez B., Rabbia Y., Nisenson P., Gezari D. Astron. Astrophys. Suppl. Ser. 134, 545 (1999) Speckle observations of double and multiple stars at Pic du Midi: measurements during 1995 and 1997 and new orbits

J/A+AS/134/561 See VI/107/

J/A+AS/135/29/  table2.dat table3.dat  (25Kb)

Hric L., Galis R. Astron. Astrophys. Suppl. Ser. 135, 29 (1999) New periods of variable stars in Cygnus based on a quarter of century of photographic observations.

J/A+AS/135/133/  redcat.dat  (4Kb)

Vauglin I., Paturel G., Borsenberger J., Fouque P., Epchtein N., Kimeswenger S., Tiphene D., Lanoix P., Di Nella-Courtois H. Astron. Astrophys. Suppl. Ser. 135, 133 (1999) First DENIS I-band extragalactic catalog

J/A+AS/135/145/  catalog.dat  (52Kb)

Rozas, M. Zurita, A., Heller, C. H., Beckman, J. E. Astron. Astrophys. Suppl. Ser. 135, 145 (1999) Global properties of the population of HII regions in NGC 7479 from photometric Hαimaging.

J/A+AS/135/159/  tab8-10.dat table11.dat table12.dat  (11Kb)

Mendoza C., Zeippen C.J., Storey P.J. Astron. Astrophys. Suppl. Ser. 135, 159 (1999) Atomic data from the IRON Project. XXXIII. Radiative rates for the intercombination transitions in the carbon isoelectronic sequence.

J/A+AS/135/227/  table.dat  (15Kb)

Jilinski E.G., Puliaev S., Penna J.L., Andrei A.H., Laclare F. Astron. Astrophys. Suppl. Ser. 135, 227 (1999) Solar diameter observations at Observatorio Nacional in 1997-1998.

J/A+AS/135/231/  table1.dat table2.dat table3.dat table4.dat table5.dat table6.dat table7.dat  (38Kb)

Falin J.L., Mignard F. Astron. Astrophys. Suppl. Ser. 135, 231 (1999) Mining in the Hipparcos raw data

J/A+AS/135/273/  table2.dat table3.dat table4.dat  (19Kb)

Marecki A., Falcke H., Niezgoda J., Garrington S.T., Patnaik A.R. Astron. Astrophys. Suppl. Ser. 135, 273 (1999) Gigahertz Peaked Spectrum sources from the Jodrell Bank-VLA Astrometric Survey I. Sources in the region 35^{∘}=δ=75^{∘}

J/A+AS/135/319/  table2.dat  (127Kb)

Huensch M., Schmitt H.H.M.M., Sterzik M.F., Voges W. Astron. Astrophys. Suppl. Ser. 135, 319 (1999) The ROSAT all-sky survey catalogue of the nearby stars

J/A+AS/135/347/  aij24.dat aij25.dat elfe24.dat elfe25.dat tarfe24.dat tarfe25.dat  (236Kb)

Nahar S N, Pradhan A.K. Astron. Astrophys. Suppl. Ser. 135, 347 (1999)

Atomic data from the Iron Project XXXV. Relativistic fine structure oscillator strengths for Fe XXIV and Fe XXV

J/A+AS/135/359/ tab34.dat table1.dat table2a.dat table2b.dat table5.dat table6.dat table7.dat (71Kb) Clegg R.E.S., Miller S., Storey P.J., Kisielius R. Astron. Astrophys. Suppl. Ser. 135, 359 (1999) Recombination line intensities for hydrogenic ions: the fine structure components of HI and HeII

J/A+AS/135/391/  table2.dat  (1.8 Mb)

Sagar R., Subramaniam A., Richtler T., Grebel R.K. Astron. Astrophys. Suppl. Ser. 135, 391 (1999) VI CCD photometry of metal-rich bulge globular clusters: NGC 6553

J/A+AS/135/405/ table1.dat table2.dat table3.dat table4.dat table5.dat table6.dat table7.dat table8.dat table9.dat table10.dat table11.dat table12.dat table13.dat table14.dat table15.dat table16.dat table17.dat table18.dat table19.dat table20.dat table21.dat (136Kb) Charbonnel C., Dappen W., Schaerer D., Bernasconi P.A., Maeder A., Meynet G., Mowlavi N. Astron. Astrophys. Suppl. Ser. 135, 405 (1999) Grids of stellar models. VIII. From 0.4 to 1.0 M_{Sun} at Z=0.020 and Z=0.001, with the MHD equation of state

J/A+AS/135/437/  table5.dat table6.dat  (24Kb)

Goncalves A.C., Veron-Cetty M.-P., Veron P. Astron. Astrophys. Suppl. Ser. 135, 437 (1999) AGNs with composite spectra. II. Additional data.

J/A+AS/135/493/  table2.dat  (14Kb)

Handler G. Astron. Astrophys. Suppl. Ser. 135, 493 (1999) Variable central stars of young Planetary Nebulae. A photometric study of the central star of M2-54

J/A+AS/135/499/  table1.dat  (8Kb)

Salaman M., Morlet G., Gili R. Astron. Astrophys. Suppl. Ser. 135, 499 (1999) CCD measurements of visual double stars made with the 50 cm refractor of the Nice Observatory.

J/A+AS/135/503/  table6.dat table7.dat table8.dat table9.dat  (157Kb)

Grenier S., Burnage R., Farraggiana R., Gerbaldi M., Delmas F., Gomez A.E., Sabas V., Sharif L. Astron. Astrophys. Suppl. Ser. 135, 503 (1999) Radial velocities of Hipparcos southern B8-F2 type stars

J/A+AS/135/511/  table1.dat table2.dat table3.dat table4.dat table5.dat tables.ps tables.tex  (140Kb)

Ugryumov A.V., Engels D., Lipovetsky V.A., Hagen H.-J., Hopp U., Pustilnik S.A., Kniazev A.Yu., Richter G., Izotov Yu.I., Popescu C.C. Astron. Astrophys. Suppl. Ser. 135, 511 (1999) The Hamburg/SAO Survey for Emission-Line Galaxies. I. A First List of 70 Galaxies.

J/A+AS/135/555/  model.dat table1.dat table2.dat table3.dat  (10.9 Mb)

Alencar S.H.P., Vaz L.P.R. Astron. Astrophys. Suppl. Ser. 135, 555 (1999) Limb-darkening coefficients of illuminated atmospheres 1. Results for illuminated line-blanketed models with 3700K Teff 7000K

J/A+AS/135/571/  table4.dat  (21Kb)

Zukowski E.L.H., Kronberg P.P., Forkert T., Wielebinski R. Astron. Astrophys. Suppl. Ser. 135, 571 (1999) Linear polarization measurements of extragalactic radio sources at λ6.3 cm.

J/A+AS/136/1/  cpasj2.dat  (529Kb)

Manrique W.T., Lu L., Perdomo R., Podesta R.C., Wang Z., Actis E.L., Zeng F., Zeng Z., Alonso E., Pacheco A.M., Bustos G. Astron. Astrophys. Suppl. Ser. 136, 1 (1999) Second San Juan photoelectric astrolabe catalogue

J/A+AS/136/35/  refs.dat table1.dat table2.dat table3.dat tables.ps tables.tex  (107Kb)

Schaerer D., Contini T., Pindao M. Astron. Astrophys. Suppl. Ser. 136, 35 (1999) New catalogue of Wolf-Rayet galaxies and high-excitation extra-galactic HII regions

J/A+AS/136/81/  table1.dat  (38Kb)

Kahabka P., Pietsch W., Filipovic M.D., Haberl F. Astron. Astrophys. Suppl. Ser. 136, 81 (1999)

A ROSAT PSPC X-Ray Survey of the Small Magellanic Cloud

J/A+AS/136/95/  table3.dat table3.ps table3.tex  (69Kb)

Rossi C., Errico L., Friedjung M., Giovannelli F., Muratorio G., Viotti R., Vittone A. Astron. Astrophys. Suppl. Ser. 136, 95 (1999) On the nature of the Herbig Be star V380 Orionis

J/A+AS/136/111/  table2.dat table3.dat table4.dat table6.dat tablea.dat  (39Kb)

Adelman S.J., Rice R.H. Astron. Astrophys. Suppl. Ser. 136, 111 (1999) uvby photometry of the magnetic CP stars HD 35298, 19 Lyrae, HD 192678, and HR 8216

J/A+AS/136/139/  table2.dat table3.dat  (22Kb)

Yang Y., Liu Q. Astron. Astrophys. Suppl. Ser. 136, 139 (1999) Photometric study of the late-type contact binary YY Eridani

J/A+AS/136/257/  table6.dat table7.dat table8.dat table9.dat  (57Kb)

Harper D., Beurle K.,I.P.Williams, Murray C.D., Taylor D.B., Fitzsimmons A., Cartwright I.M. Astron. Astrophys. Suppl. Ser. 136, 257 (1999) CCD astrometry of Saturn’s satellites in 1995 and 1997

J/A+AS/136/261/  tableb.dat  (43Kb)

Dominici T.P., Teixeira R., Horvath J.E., Medina Tanco G.A., Benevides-Soares P. Astron. Astrophys. Suppl. Ser. 136, 261 (1999) Extension of Tycho catalog for low-extinction windows in the galactic bulge

J/A+AS/136/293/  table1.dat table2.dat  (11Kb)

Jasinta D.M.D., Hidayat B. Astron. Astrophys. Suppl. Ser. 136, 293 (1999) The photographic observations on the components of Sirius

J/A+AS/136/307/  table5.dat  (219Kb)

Abad C., Vicente B. Astron. Astrophys. Suppl. Ser. 136, 307 (1999) An astrometric catalogue for the area of Coma Berenices

J/A+AS/136/373/  table4.dat  (21Kb)

Tokovinin A.A. Astron. Astrophys. Suppl. Ser. 136, 373 (1999) Orbits of new spectroscopic components in 8 multiple systems

J/A+AS/136/379/  table2.dat table3.dat table4.dat table5.dat tablea.dat  (37Kb)

Adelman S.J., Rayle K.E., Pi C.-L. M. Astron. Astrophys. Suppl. Ser. 136, 379 (1999) uvby Photometry of the magnetic CP stars 45 Leonis, HR 4330, 49 Herculis, and HR 6718

J/A+AS/136/395/  level4 prog.f table4  (703Kb)

Chen G.X., Pradhan A.K. Astron. Astrophys. Suppl. Ser. 136, 395 (1999) Atomic data from the Iron Project. XXXVII. Electron impact excitation collision strengths and rate coefficients for Fe VI

J/A+AS/136/429/  table1.dat  (37Kb)

Corporon P., Lagrange-Henri A.-M. Astron. Astrophys. Suppl. Ser. 136, 429 (1999) A search for spectroscopic binaries among Herbig Ae/Be stars

J/A+AS/136/445/  table1.dat table2.dat  (25Kb)

Veiga C., Vieira Martins R., Le Guyader C. Astron. Astrophys. Suppl. Ser. 136, 445 (1999) CCD observations of Nereid and a new orbital determination

J/A+AS/136/453/  stars.dat table2.dat table3.dat  (229Kb)

Zacs L., Alksnis O., Musaev F.A., Bikmaev I.F., Galazutdinov G.A. Astron. Astrophys. Suppl. Ser. 136, 453 (1999) An abundance analysis of the single-lined spectroscopic binaries with barium stars-like orbital elements. II. The spectroscopic data.

J/A+AS/136/455/  table1.dat  (57Kb)

Bai J.M, Xie G.Z., Li K.H., Zhang X., Liu W.W. Astron. Astrophys. Suppl. Ser. 136, 455 (1999)

Photometric monitoring of three BL Lac objects in 1993-1998

J/A+AS/136/471/  table4.dat  (94Kb)

Massi F., Giannini T., Lorenzetti D., Liseau R., Moneti A., Andreani P. Astron. Astrophys. Suppl. Ser. 136, 471 (1999) Star formation in the Vela Molecular Clouds. III. Near infrared images and mm photometry of D-cloud IRAS sources.

J/A+AS/136/509/  table1.dat table2.dat table3.dat  (36Kb)

Heraudeau, P., Simien, F., Maubon, G., Prugniel, P. Astron. Astrophys. Suppl. Ser. 136, 509 (1999) Stellar kinematical data for the central region of spiral galaxies. II.

J/A+AS/136/515/  table4.dat  (12Kb)

Li Z.P., Fang M.J. Astron. Astrophys. Suppl. Ser. 136, 515 (1999) Multi-period pulsation and amplitude variations behavior in delta Scuti variable DQ Cephei

J/A+AS/136/519/  table1.dat table3.dat table4.dat  (30Kb)

Golev V., Prugniel P., Simien F., Longhetti M. Astron. Astrophys. Suppl. Ser. 136, 519 (1999) Central Mg_{2} indices for early-type galaxies.

J/A+AS/136/525/  table3.dat table3.ps table3.tex  (49Kb)

Sinachopoulos D., Dapergolas A., van Dessel E., Kontizas E. Astron. Astrophys. Suppl. Ser. 136, 525 (1999) CCD astrometry and instrumental DeltaV photometry of visual double stars. V. Differential measurements of pairs with angular separations 1” to 4”

J/A+AS/136/531/  table1.dat tableref.dat  (61Kb)

Lopes P.A.A., Andrei A.H., Puliaev S., Penna J.L., Teixeira R., Benevides-Soares P., Assafin M., Le Campion J.F., Requieme Y. Astron. Astrophys. Suppl. Ser. 136, 531 (1999) Observations of radio stars at the Valinhos CCD Meridian Circle

J/A+AS/136/571/  table1.dat table2.dat  (24Kb)

Han J.L., Tian W.W. Astron. Astrophys. Suppl. Ser. 136, 571 (1999) Pulsars identified from the NRAO VLA Sky Survey

J/A+AS/137/43/  object.dat table3.dat  (78Kb)

Forster J.R., Caswell J.L. Astron. Astrophys. Suppl. Ser. 137, 43 (1999) OH and H2O masers in 74 star-forming regions. The FC89 database

J/A+AS/137/75/  table2.dat table3.dat table4.dat table5.dat table6.dat table7.dat table8.dat  (39Kb)

Zaggia S., Hook I., Mendez R., Da Costa L., Olsen L.F., Nonino M., Wicenec A., Benoist C., Deul E., Erben T., Guarnieri M.D., Hook R., Prandoni I., Scodeggio M., Slijkhuis R., Wichmann R. Astron. Astrophys. Suppl. Ser. 137, 75 (1999) ESO Imaging Survey. IV. Multicolour analysis of point-like objects toward the South Galactic Pole.

J/A+AS/137/83/  table1.dat table2.dat  (22Kb)

Scodeggio M., Olsen L.F., Da Costa L., Slijkhuis R., Benoist C., Deul E., Erben T., Hook R., Nonino M., Wicenec A., Zaggia S. Astron. Astrophys. Suppl. Ser. 137, 83 (1999) ESO Imaging Survey. VII. Distant cluster candidates over 12 square degrees.

J/A+AS/137/101/  fig4/* table3.dat  (4.9 Mb)

Moriondo G., Baffa C., Casertano S., Chincarini G., Gavazzi G., Giovanardi C., Hunt L.K., Pierini D., Sperandio M., Trinchieri G. Astron. Astrophys. Suppl. Ser. 137, 101 (1999) Near-infrared observations of galaxies in Pisces-Perseus: I. H-band surface photometry of 174 spirals

J/A+AS/137/113/  table4.dat table4.ps table4.tex  (516Kb)

van Houten C.J., Walraven T., Walraven J.H. Astron. Astrophys. Suppl. Ser. 137, 113 (1999) Five-colour photometry of OB-stars in the Southern Hemisphere.

J/A+AS/137/157/  table1.dat table2.dat table3.dat table4.dat table5.dat  (19Kb)

Kisielius R., Storey P.J. Astron. Astrophys. Suppl. Ser. 137, 157 (1999)

Recombination coefficients for the 5g-4f transitions of O III at nebular temperatures and densities

J/A+AS/137/165/  table5.dat table5.ps table5.tex  (205Kb)

Eissner W., Galavis M.E., Mendoza C., Zeippen Astron. Astrophys. Suppl. Ser. 137, 165 (1999) Atomic data from the IRON Project. XXXVIII. Electron impact excitation of the fine-structure transitions in the n=3 complex of Fe XV

J/A+AS/137/227/  table4.ps table4.tex table4a.dat table4b.dat  (130Kb)

Lopez-Garica Z., Adelman S.J. Astron. Astrophys. Suppl. Ser. 137, 227 (1999) Elemental abundance studies of CP stars. II. The silicon stars HD 133029 and HD 192913

J/A+AS/137/233/  table1.dat table1.ps table1.tex  (247Kb)

Baume G., Vazquez R.A., Feinstein A. Astron. Astrophys. Suppl. Ser. 137, 233 (1999) UBVI imaging photometry of NGC 6231

J/A+AS/137/269/  catalog.dat  (23Kb)

Martin V.A.F., Leister N.V., Vigouroux G., Furia M., Journet A. Astron. Astrophys. Suppl. Ser. 137, 269 (1999) Absolute declinations with the photoelectric astrolabe at Calern Observatory (OCA)

J/A+AS/137/273/  table1.dat table2.dat table3.dat table4.dat  (35Kb)

Gerbaldi M., Faraggiana R., Burnage R., Delmas F., Gomez A.E., Grenier S. Astron. Astrophys. Suppl. Ser. 137, 273 (1999) Search for reference A0 dwarf stars: masses and luminosities revisited with Hipparcos parallaxes

J/A+AS/137/293/  table2.dat table2.ps table2.tex  (1.1 Mb)

Weinberger R., Gajdosik M., Zanin C. Astron. Astrophys. Suppl. Ser. 137, 293 (1999) A visual survey for small non-stellar objects in the galactic plane on POSS II. Results of nineteen fields at 115^{∘}l 157^{∘}

J/A+AS/137/299/  table2.dat table3.dat table4.dat table5.dat table6.dat tables.ps tables.tex  (138Kb)

Pustilnik S.A., Engels D., Ugryumov A.V., Lipovetsky V.A., Hagen H.-J., Kniazev A.Yu., Izotov Yu.I., Richter G. Astron. Astrophys. Suppl. Ser. 137, 299 (1999) The Hamburg/SAO Survey for Emission-Line Galaxies. II. The Second List of 128 Galaxies

J/A+AS/137/305/  catalog.dat  (159Kb)

Scholz R.-D., Brunzendorf J., Ivanov G., Kharchenko N., Lasker B., Meusinger H., Preibisch T., Schilbach E., Zinnecker H. Astron. Astrophys. Suppl. Ser. 137, 305 (1999) IC 348 proper motion study from digitised Schmidt plates.

J/A+AS/137/351/  fig/* table1.dat tables.ps tables.tex  (9.7 Mb)

Verdugo E., Talavera A., Gomez de Castro A.I. Astron. Astrophys. Suppl. Ser. 137, 351 (1999) Understanding A-type supergiants. I. Ultraviolet and visible spectral atlas of A-type supergiants

J/A+AS/137/369/  appena.dat table5.dat  (76Kb)

Fekel C., Strassmeier K.G., Weber M., Washuettl A. Astron. Astrophys. Suppl. Ser. 137, 585 (1999) Orbital elements and physical parameters of ten chromospherically active binary stars

J/A+AS/137/419/  table6.dat  (15Kb)

Oestlin G., Amram P., Masegosa J., Bergvall N., Boulesteix J. Astron. Astrophys. Suppl. Ser. 137, 419 (1999) Dynamics of blue compact galaxies as revealed by their H-alpha velocity fields. I. The data, velocity fields and rotation curves.

J/A+AS/137/451/  table4.dat table5.dat table6.dat  (406Kb)

Grenier S., Baylac M.O., Rolland L., Burnage R., Arenou F., Briot D., Delmas F., Duflot M., Genty V., Gomez A.E., Halbwachs J.-L., Marouard M., Oblak E., Sellier A. Astron. Astrophys. Suppl. Ser. 137, 451 (1999) Radial velocities. IX. Measurements of 2800 B2-F5 stars for Hipparcos

J/A+AS/137/495/  galaxies.dat table7.dat table8.dat table9.dat  (20Kb)

Severgini P., Garilli B., Saracco P., Chincarini G. Astron. Astrophys. Suppl. Ser. 137, 495 (1999)

An Hαcatalogue of galaxies in Hickson Compact Groups. I. The Sample

J/A+AS/137/521/  table2.dat table3.dat table4.dat  (26Kb)

Munari U., Tomasella L. Astron. Astrophys. Suppl. Ser. 137, 521 (1999) High resolution spectroscopy over lambda 8500-8750Å. I. Mapping the MKK classification system.

J/A+AS/138/87/ average.dat dcurves.dat fig/* figures.dat lcurves.dat phot.dat table1.dat table2.dat table5.dat tables.ps tables.tex (397Kb) Cutispoto G., Pastori L., Tagliaferri G., Messina S., Pallavicini R., Rodono M. Astron. Astrophys. Suppl. Ser. 138, 87 (1999) Classification of EUV stellar sources detected by the ROSAT WFC I. Photometric and radial velocity studies

J/A+AS/138/247/  table1.dat  (66Kb)

Veiga C.H., Vieira Martins R. Astron. Astrophys. Suppl. Ser. 138, 247 (1999) CCD astrometric observations of Uranian satellites: 1995-1998

J/A+AS/138/253/  table2a.dat table2b.dat  (23Kb)

Boffi F.R., Sparks W.B., Macchetto F.D. Astron. Astrophys. Suppl. Ser. 138, 253 (1999) A search for candidate light echoes: photometry of supernova environments

J/A+AS/139/1/  notes.dat table1.dat table2.dat table3.dat  (51Kb)

Thuan T.X., Lipovetsky V.A., Martin J.-M., Pustilnik S.A. Astron. Astrophys. Suppl. Ser. 139, 1 (1999) HI observations of blue compact galaxies from the first and second Byurakan surveys.

J/A+AS/139/25/  table2.dat  (22Kb)

Wroblewski H., Costa E. Astron. Astrophys. Suppl. Ser. 139, 25 (1999) New high proper motion stars with declinations between -5^{∘}and -30^{∘}, and right ascensions between 13h 30m and 24h

J/A+AS/139/29/  table1.dat  (46Kb)

Gorgas J., Cardiel N., Pedraz S., Gonzalez, J.J. Astron. Astrophys. Suppl. Ser. 139, 29 (1999) Empirical calibration of the lambda 4000 A break

J/A+AS/139/47/  distance.dat eclipse.dat kccdb.dat kphot.dat occult.dat  (87Kb)

Emelianov N.V., Irsmambetova T.R., Kiseleva T.P., Tejfel V.G., Vashkovjak S.N., Glushkova E.A., Kornilov V.G., Charitonova G.A. Astron. Astrophys. Suppl. Ser. 139, 47 (1999) Photometry and position observations of Saturnian satellites during their mutual eclipses and occultations in 1995 performed at the observatories in Russia and Kazakhstan

J/A+AS/139/69/  table1.dat table2.dat table3.dat  (29Kb)

Shatsky N., Sinachopoulos D., Prado P., van Dessel E. Astron. Astrophys. Suppl. Ser. 139, 69 (1999) UBV absolute CCD photometry and differential astrometry of a sample of visual double stars with A-type primaries

J/A+AS/139/97/  appen.dat  (29Kb)

Karachentsev I.D., Makarov D.I., Huchtmeier W.K. Astron. Astrophys. Suppl. Ser. 139, 97 (1999) HI properties of nearby galaxies from a volume-limited sample

J/A+AS/139/135/  table2.dat table3.dat  (51Kb)

Crawford F.L., McKenna F.C., Keenan F.P., Aller L.H., Feibelman W.A., Ryan S.G. Astron. Astrophys. Suppl. Ser. 139, 135 (1999) Line identifications and intensities for the optical spectrum of RR Telescopii between 3180 and 9455 Å

J/A+AS/139/141/  catalog.dat  (65Kb)

Brunzendorf J., Meusinger H. Astron. Astrophys. Suppl. Ser. 139, 141 (1999) The galaxy cluster Abell 426 (Perseus). A catalogue of 660 galaxy positions, isophotal magnitudes and morphological types

J/A+AS/139/231/  table2.dat table3.dat table4.dat table5.dat table6.dat  (34Kb)

Dufton P.L., Smartt S.J., Hambly N.C. Astron. Astrophys. Suppl. Ser. 139, 231 (1999)

A UKST survey of blue objects towards the Galactic Centre - a search for early-type stars.

J/A+AS/139/277/ refs.dat table1.dat table4.dat table5.dat table6.dat table7.dat table8.dat table9.dat table10.dat (153Kb) Haberl F., Pietsch W. Astron. Astrophys. Suppl. Ser. 139, 277 (1999) A ROSAT PSPC catalogue of X-ray sources in the LMC region

J/A+AS/139/305/  table2.dat  (71Kb)

Veiga C.H., Vieira Martins R. Astron. Astrophys. Suppl. Ser. 139, 305 (1999) Photographic positions for the first eight satellites of Saturn

J/A+AS/139/321/  stars.dat  (51Kb)

Dominici T.P., Horvath J.E., Medina Tanco G.A., Teixeira R., Benevides-Soares P. Astron. Astrophys. Suppl. Ser. 139, 321 (1999) Stellar variability in low-extinction regions towards the galactic bulge

J/A+AS/139/359/  table2.dat  (112Kb)

Vigotti M., Gregorini L., Klein U., Mack K.-H. Astron. Astrophys. Suppl. Ser. 139, 1 (1999) Multi-Frequency Study of the B3-VLA Sample. II. The database

J/A+AS/139/393/  table4.dat  (46Kb)

Larsen S.S. Astron. Astrophys. Suppl. Ser. 139, 393 (1999) Young massive star clusters in nearby galaxies. II. Software tools, data reductions and cluster sizes

J/A+AS/139/425/  table1.dat table2.dat table3.dat table4.dat  (79Kb)

Cuypers J., Seggewiss W. Astron. Astrophys. Suppl. Ser. 139, 425 (1999) CCD photometry and astrometry of visual double and multiple stars of the HIPPARCOS catalogue. II. CCD photometry and differential astrometry of 288 southern ”intermediate” systems

J/A+AS/139/433/  table5.dat table5.ps table5.tex  (681Kb)

de Medeiros J.R., Mayor M. Astron. Astrophys. Suppl. Ser. 139, 433 (1999) A catalog of rotational and radial velocities for evolved stars.

J/A+AS/139/483/  table3.dat table4.dat  (14Kb)

Bonfanti, P., Simien, F., Rampazzo, R., Prugniel, P. Astron. Astrophys. Suppl. Ser. 139, 483 (1999) Kinematics of early-type galaxies in compact groups. HCG 67, HCG 74, and HCG 79

J/A+AS/139/491/  table1.dat table3.dat  (283Kb)

Makarova L. Astron. Astrophys. Suppl. Ser. 139, 491 (1999) Multi-colour photometry of nearby dwarf galaxies

J/A+AS/139/513/  table4.dat table4.ps table4.tex  (395Kb)

Jetsu L., Tuominen I., Bopp B.W., Ezhkova O.V., Grankin K.N., Hackman T., Heckert P.A., Ibragimov M.A., Kondratiev V.B., Mel’nikov S.Yu., Oja T., Nations H.L., Shevchenko V.S. Astron. Astrophys. Suppl. Ser. 139, 513 (1999) Photometry of V1794 Cygni between 1975 and 1995

J/A+AS/139/525/  table3.dat table4.dat  (48Kb)

Durret F., Felenbok P., Lobo C., Slezak E. Astron. Astrophys. Suppl. Ser. 139, 525 (1999) A catalogue of velocities in the direction of the cluster of galaxies Abell 496

J/A+AS/139/531 See II/227/

J/A+AS/139/545/  table1.dat table5.dat  (91Kb)

Kovalev Y.Y., Nizhelsky N.A., Kovalev Yu.A., Berlin A.B., Zhekanis G.V., Mingaliev M.G., Bogdantsov A.V. Astron. Astrophys. Suppl. Ser. 139, 545 (1999) Survey of instantaneous 1-22 GHz spectra of 550 compact extragalactic objects with declinations from -30^{∘}to +43^{∘}

J/A+AS/139/555/ keywords.dat refs.dat starref.dat table1.dat table2.dat table3.dat tables.tex (507Kb) Gershberg R.E., Katsova M.M., Lovkaya M.N., Terebizh A.V., Shakhovskaya N.I. Astron. Astrophys. Suppl. Ser. 139, 555 (1999) Catalogue and bibliography of UV Cet stars flare stars and related objects in the solar vicinity

J/A+AS/139/559/  table1.dat table2.dat  (353Kb)

Slezak E., Durret F., Guibert J., Lobo C. Astron. Astrophys. Suppl. Ser. 139, 559 (1999) A photometric catalogue of galaxies in the cluster Abell 496.

J/A+AS/139/575/  table2.dat  (23Kb)

Wei J.Y., Xu D.W., Dong X.Y., Hu J.Y. Astron. Astrophys. Suppl. Ser. 139, 575 (1999) An AGN sample with high X-ray-to-optical flux ratio from RASS. I. The optical identification.

J/A+AS/140/29/ amadeus.dat apt1q97.dat apt2q97.dat apt4q96.dat table1.dat wolfgang.dat (1.2 Mb) Strassmeier K.G., Serkowitsch E., Granzer T. Astron. Astrophys. Suppl. Ser. 140, 29 (1999) Starspot photometry with robotic telescopes. UBV(RI)_{c} and by light curves of 47 stars in 1996/97

J/A+AS/140/55/  appenb.dat  (22Kb)

Barthes D., Luri X., Alvarez R., Mennessier M. O. Astron. Astrophys. Suppl. Ser. 140, 55 (1999) Period-luminosity-colour distribution and classification of galactic oxygen-rich LPVs. I. Luminosity calibrations.

J/A+AS/140/79/  object.dat table1.dat table1.ps table1.tex  (199Kb)

Imbert M. Astron. Astrophys. Suppl. Ser. 140, 79 (1999) Determination of the radii of Cepheids. V. Radial velocities and dimensions of 22 galactic Cepheids (Determination des rayons de Cepheides. V. Vitesses radiales et dimensions de 22 Cepheides galactiques)

J/A+AS/140/89/  table1.dat table2.dat  (1.0 Mb)

Paturel G., Petit C., Prugniel P., Garnier R. Astron. Astrophys. Suppl. Ser. 140, 89 (1999) Galaxy coordinates. II: Accurate equatorial coordinates for 17335 galaxies

J/A+AS/140/93 See VI/98A/

J/A+AS/140/135/ ngc3680.fit ngc3680a.dat ngc3680u.dat ngc6134.fit ngc6134a.dat ngc6134u.dat (13.6 Mb) Bruntt H., Frandsen S., Kjeldsen H., Andersen M.I. Astron. Astrophys. Suppl. Ser. 140, 135 (1999) Stromgren Photometry of open clusters NGC 3680 and NGC 6134

J/A+AS/140/155/  mostra.dat  (157Kb)

Moretti A., Molinari E., Chincarini G., De Grandi S. Astron. Astrophys. Suppl. Ser. 140, 155 (1999) The luminosity function of cluster galaxies. II. Data reduction procedures applied to the cluster Abell 496.

J/A+AS/140/191/  table1.dat table2.dat table3.dat  (631Kb)

Dimitrijevic M.S., Sahal-Brechot S. Astron. Astrophys. Suppl. Ser. 140, 191 (1999) Stark broadening of neutral calcium spectral lines

J/A+AS/140/193/  table1.dat  (78Kb)

Dimitrijevic M.S., Sahal-Brechot S. Astron. Astrophys. Suppl. Ser. 140, 193 (1999) Stark broadening of neutral zinc spectral lines

J/A+AS/140/287/  table1.dat table2.dat table3.dat  (15Kb)

Balega I.I., Balega Y.Y., Maksimov A.F., Pluzhnik E.A., Shkhagosheva Z.U., Vasyuk V.A. Astron. Astrophys. Suppl. Ser. 140, 287 (1999) Binary star speckle measurements during 1992-1997 from the SAO 6-m and 1-m telescopes in Zelenchuk.

J/A+AS/140/327/  table5.dat table6.dat tables.ps tables.tex  (80Kb)

Mueller K.R., Wegner G., Raychaudhury S., Freudling W. Astron. Astrophys. Suppl. Ser. 140, 327 (1999) Fundamental plane distances to early-type field galaxies in the South Equatorial Strip. I. The spectroscopic data.

J/A+AS/140/351/  table3.dat  (32Kb)

La Franca F., Lissandrini C., Cristiani S., Miller L., Hawkins M.R.S., McGillivray H.T. Astron. Astrophys. Suppl. Ser. 140, 351 (1999) A survey of UV-excess AGNs in the South Galactic Pole

J/A+AS/141/113/  table3.dat table4.dat table5.dat  (49Kb)

Andreon S., Pello R., Davoust E., Dominguez R., Poulain P. Astron. Astrophys. Suppl. Ser. 141, 113 (2000) Extensive near-infrared (H-band) photometry in Coma

J/A+AS/141/123/  table1.dat table2.dat  (536Kb)

Kraan-Korteweg R. Astron. Astrophys. Suppl. 141, 123 (2000) A catalog of galaxies behind the Southern Milky Way. I. The Hydra/Antlia extension (l= 266^{∘}-296^{∘}).

J/A+AS/141/141 See III/238/

J/A+AS/141/175/  tables.dat tables.ps tables.tex  (79Kb)

Sen A.K., Gupta Ranjan, Ramaprakash A.N., Tandon S.N. Astron. Astrophys. Suppl. Ser. 141, 175 (2000) Imaging polarimetry of some selected dark clouds

J/A+AS/141/257/  table1.dat table2.dat  (21Kb)

Aussel H., Coia D., Mazzei P., De Zotti G., Franceschini A. Astron. Astrophys. Suppl. Ser. 141, 257 (2000) ISOCAM observations of the very deep IRAS 60 micron sample in the NEP region. I. The data.

J/A+AS/141/297/ table1.dat table2.dat table3.dat table4.dat table5.dat table6.dat table7.dat table8.dat (108Kb) Abgrall H., Roueff E., Drira I. Astron. Astrophys. Suppl. Ser. 141, 297 (2000) Total transition probability and spontaneous radiative dissociation of B, C, B’ and D states of molecular hydrogen.

J/A+AS/141/371/ hbtracks.dat isochron.dat isochsum.dat mstracks.dat sspcolor.dat surfcomp.dat (12.5 Mb) Girardi L., Bressan A., Bertelli G., Chiosi C. Astron. Astrophys. Suppl. Ser. 141, 371 (2000) Evolutionary tracks and isochrones for low- and intermediate-mass stars: From 0.15 to 7 M_{⊙}, and from Z=0.0004 to 0.03.

J/A+AS/141/409/  table1.dat table3.dat  (29Kb)

Perez-Gonzalez P.G., Zamorano J., Gallego J., Gil de Paz A. Astron. Astrophys. Suppl. Ser. 141, 409 (2000) Optical photometry of the UCM Lists I and II. I. The data.

J/A+AS/141/449/  indice1.dat indice2.dat kinema.dat photom.dat table1.dat  (151Kb)

Mehlert D., Saglia R.P., Bender R., Wegner G. Astron. Astrophys. Suppl. Ser. 141, 449 (2000) Spatially resolved spectroscopy of Coma cluster early-type galaxies. I. The Database.

J/A+AS/141/469/  notes.dat table1.dat table2.dat  (59Kb)

Huchtmeier W.K., Karachentsev I.D., Karachentseva V.E., Ehle M. Astron. Astrophys. Suppl. Ser. 141, 469 (2000) HI observations of nearby galaxies. I. The first list of the Karachentsev catalog.

J/A+AS/141/491/  table3.dat table4.dat table5.dat  (44Kb)

Chen Y.Q, Nissen P.E., Zhao G., Zhang H.W., Benoni T. Astron. Astrophys. Suppl. Ser. 141, 491 (2000) Chemical composition of 90 F and G disk dwarfs

J/A+AS/142/25/  table2.dat  (224Kb)

Morras, R., Bajaja, E., Arnal, E.M., Poppel, W.G.L. Astron. Astrophys. Suppl. Ser. 142, 25 (2000) A new survey for high velocity HI detections in the Southern Hemisphere.

J/A+AS/142/41/  table2.dat  (98Kb)

Haberl F., Filipovic M.D., Pietsch W., Kahabka P. Astron. Astrophys. Suppl. Ser. 142, 41 (2000) A ROSAT PSPC catalogue of X-ray sources in the SMC region

J/A+AS/142/65/  table1.dat table2.dat  (287Kb)

Gavazzi G., Franzetti P., Scodeggio M., Boselli A., Pierini D., Baffa C., Lisi F., Hunt L.K. Astron. Astrophys. Suppl. Ser. 142, 65 (2000) 1.65 micrometer (H-band) surface photometry of galaxies. III. Observations of 558 galaxies with the TIRGO 1.5m telescope.

J/A+AS/142/73/  table1.dat table2.dat  (116Kb)

Boselli A., Gavazzi G., Franzetti P., Pierini D., Scodeggio M. Astron. Astrophys. Suppl. Ser. 142, 73 (2000) 1.65 micrometer (H-band) surface photometry of galaxies. IV: Observations of 170 galaxies with the Calar Alto 2.2m telescope.

J/A+AS/142/85/  table1.dat table2.dat table3.dat table4.dat table5.dat  (70Kb)

Davey A.R., Storey P.J., Kisielius R. Astron. Astrophys. Suppl. Ser. 142, 85 (2000) Recombination coefficients for C II

J/A+AS/142/113/  tables.dat  (788Kb)

Chandra S., Kegel W.H. Astron. Astrophys. Suppl. Ser. 142, 113 (2000) Collisional rates for asymmetrical top molecules

J/A+AS/142/181/ table1.dat table2.dat table3.dat table4.dat table5.dat table6.dat table7.dat table8.dat table9.dat tables.tex (188Kb) Cernicharo J., Guelin M., Kahane C. Astron. Astrophys. Suppl. Ser. 142, 181 (2000) A λ2mm molecular line survey of the C-star envelope IRC+10 216

J/A+AS/142/217 See III/213/

J/A+AS/142/275/  tablea1.dat tablea2.dat tablea3.dat  (276Kb)

Strassmeier K. G., Washuettl A., Granzer T., Scheck M., Weber M. Astron. Astrophys. Suppl. Ser. 142, 275 (2000) The Vienna-KPNO search for Doppler-imaging candidate stars I. A catalog of stellar-activity indicators for 1058 late-type Hipparcos stars

J/A+AS/142/347/  table2.dat  (21Kb)

Drozdovsky I., Tikhonov N. Astron. Astrophys. Suppl. Ser. 142, 347 (2000) The stellar content and distance to the nearby blue compact dwarf galaxy NGC 6789.

J/A+AS/142/353/  fig2/* prof/* table1.dat  (2.3 Mb)

Gonzalez-Serrano J.I., Carballo R. Astron. Astrophys. Suppl. Ser. 142, 353 (2000) Surface photometry of radio loud elliptical galaxies from the B2 sample.

J/A+AS/142/369/  table1.dat  (7Kb)

Wroblewski H., Costa E. Astron. Astrophys. Suppl. Ser. 142, 369 (2000) Proper motions of Luyten Catalogue stars with declinations between -5^{∘}and -30^{∘}and right ascensions between 13h 30m and 24h.

J/A+AS/142/417/  table2.dat table3.dat table4.dat table5.dat table6.dat tables.ps tables.tex  (100Kb)

Hopp U., Engels D., Green R.F., Ugryumov A.V., Izotov Y.I., Hagen H.-J., Kniazev A.Y., Lipovetsky V.A., Pustilnik S.A., Brosch N., Masegosa J., Martin J.-M., Marquez I. Astron. Astrophys. Suppl. Ser. 142, 417 (2000) The Hamburg/SAO survey for emission-line galaxies. III. The third list of 81 galaxies.

J/A+AS/142/425/  table31.dat table32.dat table33.dat table34.dat table35.dat table36.dat  (59Kb)

Drozdovsky I.O., Karachentsev I.D. Astron. Astrophys. Suppl. Ser. 142, 425 (2000) Photometric distances to six bright resolved galaxies.

J/A+AS/142/451/  table3.dat table4.dat  (12Kb)

Lamy H., Hutsemekers D. Astron. Astrophys. Suppl. Ser. 142, 451 (2000) Optical Polarization of 47 quasi-stellar objects : the data

J/A+AS/142/467/  table1.dat  (564Kb)

Barklem P.S., Piskunov N., O’Mara B.J. Astron. Astrophys. Suppl. Ser. 142, 467 (2000) A list of data for the broadening of metallic lines by neutral hydrogen collisions.

J/A+AS/143/211/  table1.dat  (90Kb)

Berdnikov L.N., Dambis A.K., Vozyakova O.V. Astron. Astrophys. Suppl. Ser. 143, 211 (2000) Galactic Cepheids. Catalogue of light-curve parameters and distances

J/A+AS/143/265/  table.dat  (11Kb)

Puliaev S., Penna J.L., Jilinski E.G., Andrei A.H. Astron. Astrophys. Suppl. Ser. 143, 265 (2000) Solar diameter observations at Observatorio Nacional in 1998-1999.

J/A+AS/143/269/  table1.dat table2.dat  (49Kb)

Szymczak M., Hrynek G., Kus A.J. Astron. Astrophys. Suppl. Ser. 143, 269 (2000) A survey of the 6.7 GHz methanol maser emission from IRAS sources. I. Data.

J/A+AS/143/303/  fitfiles.dat fits/* notes.dat tablea1.dat tablea2.dat tablea3.dat  (455.0 Mb)

De Breuck C., van Breugel W., Rottgering H.J.A., Miley G. Astron. Astrophys. Suppl. Ser. 143, 303 (2000) A sample of 669 ultra steep spectrum radio sources to find high redshift radio galaxies

J/A+AS/143/335/  table1.dat table2.dat tables.ps tables.tex  (119Kb)

Ling J.F., Prieto C. Astron. Astrophys. Suppl. Ser. 143, 335 (2000) Micrometer measurements of double stars made at the Cote D’Azur and Calar Alto observatories.

J/A+AS/143/343/  table2.dat table2.ps table2.tex  (139Kb)

Henden A., Munari U. Astron. Astrophys. Suppl. Ser. 143, 343 (2000) UBV(RI)c photometric comparison sequences for symbiotic stars

J/A+AS/143/357/  fig/* table1.dat table2.dat table3.dat  (904Kb)

Katajainen S., Takalo L.O., Sillanpaa A., Nilsson K., Pursimo T., Hanski M., Heinamaki P., Kotoneva E., Lainela M., Nurmi P., Pietila H., Rekola R., Riehokainen A., Teerikorpi P., Valtaoja E., Lahteenmaki A. Astron. Astrophys. Suppl. Ser. 143, 357 (2000) Tuorla quasar monitoring. I. Observations of 1995 - 1997

J/A+AS/143/391/  refs.dat table1.dat table2.dat table3.dat table4.dat  (71Kb)

Sasaki M., Haberl F., Pietsch W. Astron. Astrophys. Suppl. Ser. 143, 391 (2000) ROSAT HRI catalogue of X-ray sources in the LMC region

J/A+AS/143/405/  table1.dat  (5Kb)

Veiga C.H., Vieira Martins R. Astron. Astrophys. Suppl. Ser. 143, 405 (2000) Astrometric photographic observations of Helene.

J/A+AS/143/409/  appendix.dat table5.dat  (249Kb)

Foster D.C., Theissen A., Butler C.J., Rolleston W.R.J., Byrne P.B., Hawley S.L. Astron. Astrophys. Suppl. Ser. 143, 409 (2000) CCD photometry and proper motions of late-type stars in the young open cluster Stock 2.

J/A+AS/143/465/  appen.dat  (28Kb)

Hawkins M.R.S. Astron. Astrophys. Suppl. Ser. 143, 465 (2000) Quasar variability: Correlations with amplitude.

J/A+AS/143/483/  osc.dat readups.f table2.dat table4.dat ups.dat  (713Kb)

Butler K., Zeippen C.J. Astron. Astrophys. Suppl. Ser. 143, 483 (2000) Atomic data from the IRON Project. XLII: Electron impact excitation of Fe XXI.

J/A+AS/144/1 See III/204/

J/A+AS/144/5/  table3.dat table4.dat table5.dat  (20Kb)

Rosenberg A., Piotto G., Saviane I., Aparicio A. Astron. Astrophys. Suppl. Ser. 144, 5 (2000) Photometric catalog of nearby globular clusters. I. A large homogeneous (V,I) color-magnitude diagram database.

J/A+AS/144/45/  table3.dat table3.ps  (153Kb)

Fabricius C., Makarov V.V. Astron. Astrophys. Suppl. Ser. 144, 45 (2000) Hipparcos astrometry for 257 stars using Tycho-2 data

J/A+AS/144/123/  table3.dat table4.dat table5.dat table6.dat tables.ps tables.tex  (68Kb)

Wouterloot J.G.A., Heithausen A., Schreiber W., Winnewisser G Astron. Astrophys. Suppl. Ser. 144, 123 (2000) Multiline CO observations of MBM 32

J/A+AS/144/141/  table1a.dat table1b.dat table3.dat table5a.dat table5b.dat table7.dat  (98.1 Mb)

Nahar S.N., Delahaye F., Pradhan A.K., Zeippen C.J. Astron. Astrophys. Suppl. Ser. 144, 41 (2000) Atomic data from the Iron Project. XLIII. Transition probabilities for Fe V

J/A+AS/144/195/  object.dat table1.dat table1.ps table1.tex  (80Kb)

Orsatti A.M., Vega E.I., Marraco H.G. Astron. Astrophys. Suppl. Ser. 144, 195 (2000) Dust properties in NGC 6611

J/A+AS/144/227/  table2.dat  (286Kb)

Geffert M., Maintz G. Astron. Astrophys. Suppl. Ser. 144, 227 (2000) First results of a photometric and astrometric study of the globular cluster M 71 (NGC 6838).

J/A+AS/144/235 See II/228A/

J/A+AS/144/441/  fig1-9.ps fit/* fits.dat table1.dat table2.dat tables.ps tables.tex  (7.0 Mb)

Nilsson A., Bergman P., Hjalmarson A Astron. Astrophys. Suppl. Ser., 144, 441 (2000) SO and CS observations of molecular clouds I. The observational data

J/A+AS/144/469/  notes.dat refs.dat table1.dat table2.dat tables.tex  (622Kb)

Rodriguez E., Lopez-Gonzalez M.J., Lopez de Coca P. Astron. Astrophys. Suppl. Ser. 144, 469 (2000) A revised catalogue of Delta Sct stars

J/A+AS/144/475/  table1.dat table2.dat  (866Kb)

Paturel G., Petit C., Garnier R., Prugniel P. Astron. Astrophys. Suppl. Ser. 144, 475 (2000) Galaxy coordinates. III. Accurate positions for 17124 galaxies including 3301 new companions of UGC galaxies

J/A+AS/144/481/  tab2-6.dat  (11Kb)

Villata M., Raiteri C.M., Popescu M.D., Sobrito G., De Francesco G., Lanteri L., Ostorero L. Astron. Astrophys. Suppl. Ser. 144, 481 (2000) Optical observations of five X-ray selected BL Lacertae objects

J/A+AS/145/11/  table2.dat table2.ps table2.tex  (271Kb)

Melchior A.-L., Hughes S.M.G., Guibert J. Astron. Astrophys. Suppl. Ser. 145, 11 (2000) AGAPEROS: Searching for variable stars in the LMC Bar with the Pixel Method. I. Detection, astrometry and cross-identification.

J/A+AS/145/67/  table1.dat  (10Kb)

Morlet G., Salaman M., Gili R. Astron. Astrophys. Suppl. Ser. 145, 67 (2000) CCD measurements of visual double stars made with the 50 cm refractor of the Nice Observatory (2nd series)

J/A+AS/145/111/  colour.dat struct.dat ubvimag.dat  (1.6 Mb)

Volonteri M., Saracco P., Chincarini G. Astron. Astrophys. Suppl. Ser. 145, 111 (2000) A catalogue of galaxies in the HDF-South: photometry and structural parameters

J/A+AS/145/187/  table1.dat table2.dat table3.dat table4.dat  (15Kb)

Chavarria-K C., Terranegra L., Moreno-Corral M.A., de Lara E. Astron. Astrophys. Suppl. Ser. 145, 187 (2000) Photometric observations of weak-line T Tauri stars. II. WTTS in Taurus-Auriga, Orion and Scorpius OB2-2.

J/A+AS/145/223/  table3.dat  (37Kb)

Chen L., Geffert M., Wang J.J., Reif K., Braun J.M. Astron. Astrophys. Suppl. 145, 223 (2000) A proper motion study of the globular cluster M10.

J/A+AS/145/263/  table1.dat table2.dat table3.dat  (33Kb)

Simien F., Prugniel P. Astron. Astrophys. Suppl. Ser. 145, 263 (2000) Kinematical data on early-type galaxies. V.

J/A+AS/145/275/  stars.dat tables.dat  (365Kb)

Nitschelm C., Lecavelier des Etangs A., Vidal-Madjar A., Ferlet R., Olsen E.H. Astron. Astrophys. Suppl. Ser. 145, 275 (2000) A three-years Stroemgren photometric survey of suspected beta Pictoris-like stars

J/A+AS/145/365/  table1.dat table1.ps table1.tex  (222Kb)

Gupta A.C., Subramaniam A., Sagar R., Griffiths W.K. Astron. Astrophys. Suppl. Ser. 145, 365 (2000) A complete photometric study of the open cluster NGC 7790 containing Cepheid variables.

J/A+AS/145/399/  table2.dat table3.dat  (56Kb)

Bica E., Ortolani S., Barbuy B. Astron. Astrophys. Suppl. Ser. 145, 399 (2000) B and V photometry of the faint open cluster UKS 2 (BH 66)

J/A+AS/145/405/  table6.dat table7.dat tables.ps tables.tex  (395Kb)

Luetticke R., Dettmar R.-J., Pohlen M. Astron. Astrophys. Suppl. Ser. 145, 405 (2000) Box- and peanut-shaped bulges. I. Statistics.

J/A+AS/145/451/  photom/* table1.dat table2.dat table3.dat  (29.2 Mb)

Rosenberg A., Aparicio A., Saviane I., Piotto G. Astron. Astrophys. Suppl. Ser., 145, 451-465 (2000) Photometric catalog of nearby globular clusters. II. A large homogeneous (V,I) color-magnitude diagram data-base.

J/A+AS/145/491/  c1p.dat c2p.dat d1p.dat files/* ident.dat p2p.dat readme.ori  (14.7 Mb)

Gonzalez M.A., Gigosos M.A. Astron. Astrophys. Suppl. Ser. 145, 491 (2000) Tables of dipolar emission and two-photon absorption Lyman-αprofiles

J/A+AS/146/13/  table2.dat table3.dat table4.dat table5.dat tablea.dat  (40Kb)

Adelman S.J. Astron. Astrophys. Suppl. Ser. 146, 13 (2000) uvby Photometry of the magnetic CP stars 20 Eri, HR 1297, 12 CMa, and HD 134214

J/A+AS/146/41 See VIII/63/

J/A+AS/146/73/  table5.dat  (12Kb)

Tancredi G., Fernandez J.A., Rickman H., Licandro J. Astron. Astrophys. Suppl. Ser. 146, 73 (2000) A catalog of observed nuclear magnitudes of Jupiter family comets.

J/A+AS/146/103/  stars.dat table3.dat table5.dat table6.dat table7.dat  (32Kb)

Montes D., Fernandez-Figueroa M.J., De Castro E., Cornide M., Latorre A., Sanz-Forcada J. Astron. Astrophys. Suppl. Ser. 146, 103 (2000) Multiwavelength optical observations of chromospherically active binary systems. III. High resolution echelle spectra from CaII H + K to CaII IRT.

J/A+AS/146/169/  notes.dat table4.dat  (90Kb)

Galadi-Enriquez D., Trullols E., Jordi C. Astron. Astrophys. Suppl. Ser. 146, 169 (2000) Secondary UBVRI-CCD standard stars in the neighbourhood of Landolt standard stars

J/A+AS/146/217/  ik/* iklog.dat ikrefs.dat mbol.dat stars.dat vi/* vilog.dat vk/*  (5.6 Mb)

Lancon A., Wood P.R. Astron. Astrophys. Suppl. Ser. 146, 217 (2000) A library of 0.5 to 2.5 micron spectra of luminous cool stars

J/A+AS/146/251/  notes.dat objects.dat table1.dat table2.dat  (113Kb)

Baumgardt H., Dettbarn C., Wielen R. Astron. Astrophys. Suppl. Ser. 146, 251 (2000) Absolute proper motions of open clusters. I. Observational data.

J/A+AS/146/323/  table2.dat table3.dat table4.dat table5.dat  (48Kb)

Neuhaeuser R., Walter F.M., Covino E., Alcala J.M., Wolk S.J., Frink S., Guillout P., Sterzik M.F., Comeron F. Astron. Astrophys. Suppl. Ser. 146, 323 (2000) Search for young stars among ROSAT All-Sky Survey X-ray sources in and around the R CrA dark cloud

J/A+AS/146/365/  table1.dat table2.dat  (40Kb)

Kaltcheva N.T., Olsen E.H., Clausen J.V. Astron. Astrophys. Suppl. Ser. 146, 365 (2000) Stroemgren and Hβphotometry of O and B type stars in star-forming regions. III. Carina Spiral Feature

J/A+AS/146/373/  tables.dat tables.ps tables.tex  (446Kb)

Flin F., Krywult J, Trevese D., Cirimele G., Hickson P.

Astron. Astrophys. Suppl. Ser. 146, 373 (2000) Properties of nearby clusters of galaxies. IV: A147, A260, A272, A278, A1661, A2056, A2073, A2093, A2096, A2124

J/A+AS/146/407/  notes.dat refs.dat table1.dat table1b.dat table2.dat table2b.dat table8.dat  (133Kb)

Belczynski K., Mikolajewska J., Munari U., Ivison R.J., Friedjung M. Astron. Astrophys. Suppl. Ser. 146, 407 (2000) A catalogue of symbiotic stars.

J/A+AS/146/437/  table1.dat table2.dat table3.dat table4.dat table5.dat  (29Kb)

Speck A.K., Barlow M.J., Sylvester R.J., Hofmeister A.M. Astron. Astrophys. Suppl. Ser. 146, 437 (2000) Dust features in the 10-μm infrared spectra of oxygen-rich evolved stars.

J/A+AS/147/5/  table9.dat table9.ps table9.tex  (73Kb)

Miroshnichenko A.S., Chentsov E.L., Klochkova V.G., Kuratov K.S., Sheikina T.A., Mukanov D.B., Bjorkman K.S., Gray R.O., Rudy R.J., Lynch D.K., Mazuk S., Puetter R., Garcia-Lario P., Perea J.V., Bergner Yu.K. Astron. Astrophys. Suppl. Ser. 147, 5 (2000) Spectroscopy and photometry of the emission-line B-type stars AS 78 and MWC 657

J/A+AS/147/25 See V/106/

J/A+AS/147/75/  table3.dat  (20Kb)

Sasaki M., Haberl F., Pietsch W. Astron. Astrophys. Suppl. Ser. 147, 75 (2000) ROSAT HRI catalogue of X-ray sources in the SMC region.

J/A+AS/147/111/  table1.dat table2.dat table3.dat table4.dat table6.dat tables.ps  (241Kb)

Chen G.-X., Pradhan A.K. Astron. Astrophys. Suppl. Ser. 147, 111 (2000) Iron Project. XLIV. Transition probabilities and line ratios for Fe VI with fluorescent excitation in planetary nebulae

J/A+AS/147/169/  table1.dat table1.ps table1.tex table2.dat table2.ps table2.tex  (898Kb)

Bicay M.D., Stepanian J.A., Chavushyan V.H., Erastova L.K., Ayvazyan V.T., Seal J., Kojoian G. Astron. Astrophys. Suppl. Ser. 147, 169 (2000) Accurate optical positions for 2978 objects from the second Byurakan Survey (SBS) with the digitized sky survey

J/A+AS/147/187/  table1.dat table2.dat table2.ps table2.tex  (52Kb)

Huchtmeier W., Karachentsev I.D., Karachentseva V.E. Astron. Astrophys. Suppl. Ser. 147, 187 (2000) HI observations of nearby galaxies. II. The second list of the Karachentsev catalog

J/A+AS/147/195/  table2.dat  (23Kb)

Maron O., Kijak J., Kramer M., Wielebinski R. Astron. Astrophys. Suppl. Ser. 147, 195 (2000) Pulsar spectra of radio emission.

J/A+AS/147/251/  table1.dat  (8Kb)

Thorel J.-C. Astron. Astrophys. Suppl. Ser. 147, 251 (2000) Nice observatory measurements of double stars (3rd series).

J/A+AS/147/253/ table2a.dat table2b.dat table3a.dat table3b.dat table7a.dat table7b.dat table8a.dat table8b.dat tranar13.dat tranfe21.dat (33.6 Mb) Nahar S.N. Astron. Astrophys. Suppl. Ser. 147, 253 (2000) Atomic data from the Iron Project XLV. Relativistic transition probabilities for carbon-like Ar XIII and Fe XXI using Breit-Paili R-matrix method

J/AJ/61/201 See III/13/

J/AJ/63/211 See III/13/

J/AJ/67/37 See III/10A/

J/AJ/67/491 See V/6/

J/AJ/68/194 See III/9/

J/AJ/71/990 See VII/21/ J/AJ/73/246 See III/247/

J/AJ/74/877 See III/247/

J/AJ/74/920 See II/63A/

J/AJ/75/41 See II/46/

J/AJ/75/785 See III/45/

J/AJ/77/405/  notes.dat refs.dat table2.dat table3.dat table4.dat table5.dat  (113Kb)

Bridle A.H., Davis M.M., Fomalont E.B., Lequeux J. Astron. J. 77, 405 (1972) Flux densities, positions, and structures for a complete sample of intense radio sources at 1400MHz.

J/AJ/77/797/  notes.dat table2.dat  (28Kb)

Pauliny-Toth I.I.K., Kellermann K.I. Astron. J. 77, 797 (1972) The NRAO 5-GHz radio source survey. III. The 140-ft ”strong” source survey.

J/AJ/78/687 See III/46/

J/AJ/78/959 See VI/19/

J/AJ/79/1356 See VII/4A/

J/AJ/80/208 See VII/218/

J/AJ/80/379/  table3  (67Kb)

HANSON R.B. Astron. J. 80, 379 (1975) A study of the motion, membership, and distance of the Hyades cluster

J/AJ/80/771/  notes.dat table2.dat  (15Kb)

Sramek R. Astron. J. 80, 771 (1975) 5-GHz survey of bright galaxies.

J/AJ/80/972/  table.dat  (16Kb)

Stephenson C.B., Sanduleak N. Astron. J., 80, 972-975 (1975) Dwarf K and M stars discovered on objective-prism plates.

J/AJ/80/1011/  table2  (52Kb)

HANSEN O.L., BLANCO V.M. Astron. J. 80, 1011 (1975) Classification of 831 two-micron sky survey sources south of +5deg

J/AJ/81/807/  table2.dat  (22Kb)

Young P.J. Astron. J. 81, 807 (1976) Tables of functions for a spherical galaxy obeying the r^{1/4} law in projection

J/AJ/82/113 See II/149/

J/AJ/83/1549/  table3  (43Kb)

Kirshner R.P., Oemler A.Jr., Schechter P.L. Astron. J. 83, 1549 (1978) A study of field galaxies. I. Redshifts and photometry of a complete sample of galaxies

J/AJ/84/783 See VI/19/

J/AJ/84/1305/  table1.dat table2.dat table3.dat  (10Kb)

van den Bergh S., Younger F. Astron. J. 84, 1305 (1979) The metal-rich globular cluster NGC 6528.

J/AJ/87/739/  table1 table2  (26Kb)

Condon J.J., Condon M.A., Hazard C. Astron. J. 87, 739 (1982) Complete samples of active extragalactic objects. I. A 1411-MHz VLA survey centered on alpha= 12h04min, delta= +11deg30min

J/AJ/87/859/  notes.dat table2.dat  (75Kb)

Perley R.A. Astron. J. 87, 859 (1982) The positions, structures and polarizations of 404 compact radio sources.

J/AJ/87/1083 See VIII/29/

J/AJ/87/1165/ asc2bin.for extinctn.for hinorth.asc hisouth.asc north.gif poles.gif redexpla.doc rednorth.asc redsouth.asc south.gif (2.8 Mb) Burstein A., Heiles C. Astron. J. 87, 1165 (1982) Reddening Maps derived from HI and Galaxy Counts

J/AJ/87/1429/  table1  (10Kb)

Condon J.J., Mitchell K.J. Astron. J. 87, 1429 (1982) Complete samples of active extragalactic objects. II. A deep 1.452-GHz VLA survey centered on 08:52:15+17:16

J/AJ/87/1515/  table6.dat  (22Kb)

Pier J.R. Astron. J. 87, 1515 (1982) Stars of spectral type A and B in the southern galactic halo. I. UBV photometry

J/AJ/87/1679 See V/33/

J/AJ/88/439 See II/118/

J/AJ/88/1285/  table2  (28Kb)

Kirshner R.P., Oemler A., Schechter P.L., Shectman S.A. Astron. J. 88, 1285 (1983) A deep survey of galaxies

J/AJ/88/1378 See I/96/

J/AJ/89/53/  table3.dat table4.dat  (18Kb)

Sadler E.M. Astron. J. 89, 53 (1984) Radio and optical observations of a complete sample of E and SO galaxies. III. A radio continuum survey at 2.7 and 5.0 GHz.

J/AJ/89/636/  table1.dat  (20Kb)

Blanco V.M., Mccarthy M.F., Blanco B.M. Astron. J., 89, 636-647 (1984) Giant M stars in Baade’s Window.

J/AJ/89/1897/  table1.dat table2.dat table3.dat  (29Kb)

Stock J., MacConnell D.J., Osborn W., Alvarez H. Astron. J. 89, 1897 (1984) Southern stars of high radial velocity.

J/AJ/90/756 See V/48/

J/AJ/90/987 See VII/143/

J/AJ/90/1437/  table2.dat table4.dat  (28Kb)

Coleman P.H., Condon J.J., Hazard C. Astron. J. 90, 1437 (1985) Complete samples of active extragalactic objects. III. A 1411 MHz VLA survey.

J/AJ/90/1681/  vcc.dat  (148Kb)

Binggeli B., Sandage A., Tammann G.A. Astron. J. 90, 1681 (1985) Studies of the Virgo cluster. II. A catalog of 2096 galaxies in the Virgo cluster area.

J/AJ/90/2089/  table1.dat table3.dat table4.dat table5.dat  (33Kb)

Beers T.C., Preston G.W., Shectman S.A. Astron. J. 90, 2089 (1985) A search for stars of very low metal abudance. I

J/AJ/90/2221/  table4.dat  (74Kb)

Olszewski E.W., Aaronson M. Astron. J. 90, 2221 (1985) The Ursa Minor dwarf galaxy : still an old stellar system.

J/AJ/90/2540 See VII/97/

J/AJ/91/144 See III/123B/

J/AJ/91/1051 See VII/97/ J/AJ/91/1189 See III/144/

J/AJ/92/787/  notes.dat table1.dat  (41Kb)

Fich M. Astron. J. 92, 787 (1986) A complete VLA survey in the outer galaxy.

J/AJ/93/592 See III/145/

J/AJ/93/923 See III/145/

J/AJ/94/501 See I/199A/

J/AJ/95/284/  cluster.dat table1.dat  (144Kb)

Dressler A., Shectman S.A. Astron. J. 95, 284 (1988) A catalog of 1268 radial velocities in 15 rich clusters.

J/AJ/95/1629/  table1 table2  (50Kb)

Moody J.W., Kirshner R.P. Astron. J. 95, 1629 (1988) [O III] emission in two magnitude-limited field-galaxy surveys

J/AJ/96/816 See VII/143/

J/AJ/96/1076 See III/143/

J/AJ/96/1632/  table1  (22Kb)

Schade D., Vandenberg D.A., Hartwick F.D.A. Astron. J. 96, 1632 (1988) A color-magnitude diagram for the halo globular cluster M 55

J/AJ/96/1655/  table2  (35Kb)

Garwood R.W., Perley R.A., Dickey J.M., Murray M.A. Astron. J. 96, 1655 (1988) A VLA snapshot continuum survey of the first quadrant of the galactic plane at 1.5 GHz.

J/AJ/96/1884/  table1  (50Kb)

Humphreys R.M., Pennington R.L., Jones T.J., Ghigo F.D. Astron. J. 96, 1884 (1988) The red supergiants in M31: spectra, colors, and luminosities

J/AJ/97/1360/  table3  (77Kb)

Stetson P.B., Vandenberg D.A., Bolte M., Hesser J.E., Smith G.H. Astron. J. 97, 1360 (1989) CCD photometry of the anomalous globular cluster Palomar 12

J/AJ/98/64/  cluster.dat notes.dat optical.dat radio.dat refs.dat  (64Kb)

Zhao J.-H., Burns J.O., Owen F.N. Astron. J. 98, 64 (1989) A 20cm VLA survey of Abell clusters of galaxies. I. Distance class 3 clusters.

J/AJ/98/367 See VII/160/

J/AJ/98/419/  b3id.dat refs.dat  (136Kb)

Vigotti M., Grueff G., Perley R., Clark B.G., Bridle A.H. Astron. J. 98, 419 (1989) Structures, spectral indexes, and optical identifications of radio sources selected from the B3 catalogue.

J/AJ/98/1148/  /3clustrs.dat  (23Kb)

Moffet A.T., Birkinshaw M. Astron. J. 98, 1148 (1989) A VLA survey of the 3 clusters of galaxies 0016+16, Abell 665, and Abell 2218.

J/AJ/98/1305/  posmk.dat table2.dat table5.dat table7.dat  (57Kb)

Massey P., Parker J.W., Garmany C.D. Astron. J. 98, 1305 (1989) The stellar content of NGC 346: a plethora of O stars in the SMC

J/AJ/98/1693/  notes table3 table4  (205Kb)

Ryan S.G. Astron. J. 98, 1693 (1989) Subdwarf studies. I. UBVRI photometry of NLTT stars.

J/AJ/98/2086/  table6 table9 table11  (25Kb)

Walker A.R. Astron. J. 98, 2086 (1989) CCD photometry of the RR Lyrae variables in the LMC cluster NGC 2257 and the adjacent field

J/AJ/99/784/  table2 table3 table4 table5 table6  (67Kb)

Hughes S.M.G., Wood P.R. Astron. J. 99, 784 (1990) Long-period variables in the Large Magellanic Cloud. II. Infrared photometry, spectral classification, AGB evolution, and spatial distribution.

J/AJ/99/1071/  table1.dat table2.dat table3.dat table4.dat  (40Kb)

Condon J.J., Dickey J.M., Salpeter E.E. Astron. J. 99, 1071 (1990) A 1.4 GHz source survey in an area without nearby rich galaxy clusters

J/AJ/99/1079/  table2.dat table3.dat table4.dat  (12Kb)

Ciardullo R., Ford H.C., Williams R.E., Tamblyn P., Jacoby G.H. Astron. J. 99, 1079 (1990) The nova rate in the elliptical component of NGC 5128.

J/AJ/99/1284 See I/151/

J/AJ/99/1831/  table6.dat  (39Kb)

Alcaino G., Liller W., Alvarado F., Wenderoth E. Astron. J. 99, 1831 (1990) BVRI CCD photometry of the metal-poor globular cluster M68 (NGC 4590)

J/AJ/99/2019 See I/220/

J/AJ/99/2059 See I/220/

J/AJ/99/2082 See I/220/

J/AJ/100/1 See VII/180/

J/AJ/100/47 See VII/143/

J/AJ/100/1091/  table9  (356Kb)

PELETIER R.F., DAVIES R.L., ILLINGWORTH G.D., DAVIS L.E., CAWSON M. Astron. J. 100, 1091 (1990) CCD Surface Photometry of galaxies with dynamical data. II. UBR photometry of 39 elliptical galaxies

J/AJ/100/1712 See I/157/

J/AJ/100/1811/  table2 table3  (86Kb)

BUONANNO R., BUSCEMA G., FUSI PESCI F., RICHER H.B., FAHLMAN G.G. Astron. J. 100, 1811 (1990) Ruprecht 106: a young metal-poor galactic globular cluster.

J/AJ/101/1408/  mt91.dat  (52Kb)

Massey P., Thompson A.B. Astron. J. 101, 1408 (1991) Massive stars in Cyg OB2.

J/AJ/101/1561/  macro.tex table1 tb1.tex  (37Kb)

PORTER A.C., SCHNEIDER D.P., HOESSEL J.G. Astron. J. 101, 1561 (1991) CCD observations of Abell clusters. V. Isophotometry of 175 brightest elliptical galaxies in Abell clusters

J/AJ/101/1763 See II/162/

J/AJ/101/2229 See V/81/

J/AJ/102/137/  table2  (46Kb)

BENCIVENNI D., BROCATO E., BUONANNO R., CASTELLANI V. Astron. J. 102, 137 (1991) The young Magellanic cluster NGC 2004

J/AJ/102/395/  table4  (48Kb)

HUMPHREYS R.M., LANDAU R., GHIGO F.D., ZUMACH W., LABONTE A.E. Astron. J. 102, 395 (1991) BVR photoelectric sequences for selected fields in the Palomar Sky Survey and the magnitude-diameter relation

J/AJ/102/1663/  rad iras.dat  (32Kb)

Condon J.J., and Broderick J.J. Astron. J. 102, 1663 (1991)

Radio properties of extragalactic IRAS sources

J/AJ/103/267/  tables4.dat  (107Kb)

Beers T.C., Preston G.W., Shectman S.A., Doinidis S.P., Griffin K.E. Astron. J. 103, 267 (1992) Spectrocopy of hot stars in the galactic halo

J/AJ/103/460/  table2  (92Kb)

HODDER P.J.C., NEMEC J.M., RICHER H.B., FAHLMAN G.G. Astron. J. 103, 460 (1992) Deep CCD photometry and variable stars in the metal-rich globular cluster M 71

J/AJ/103/488 See V/80/

J/AJ/103/514/  table1 table34  (21Kb)

REID N., MURRAY C.A. Astron. J. 103, 514 (1992) High-velocity stars toward the South Galactic Cap.

J/AJ/103/638/  notes.dat table1.dat  (33Kb)

Monet D.G., Dahn C.C., Vrba F.J., Harris H.C., Pier J.R., Luginbuhl C.B., Ables H.D. Astron. J. 103, 638 (1992) US Naval Observatory CCD parallaxes of faint stars. I. Program description and first results.

J/AJ/103/983/  table1  (25Kb)

Bosh A.S., McDonald S.W. Astron. J. 103, 983 (1992) Stellar occultation candidates from the Guide Star Catalog. I. Saturn, 1991-1999

J/AJ/103/1205/  stars.dat ubv.dat  (231Kb)

Parker J.WM., Garmany C.D., Massey P., Walborn N.R. Astron. J. 103, 1205 (1992) The Stellar Content of LH 9 and 10 (N 11) in the LMC: A Case for Sequential Star Formation

J/AJ/103/1689 See I/176/

J/AJ/103/1987/  table5.dat  (137Kb)

Beers T.C., Preston G.W., Shectman S.A. Astron. J. 103, 1987 (1992) A search for stars of very low metal abundance. II.

J/AJ/104/28/  refs.dat table1.dat table2.dat table3.dat  (24Kb)

Bersanelli M., Bouchet P., Falomo R., Tanzi E.G. Astron. J. 104, 28 (1992) Homogeneous JHKL photometry of a sample of BL Lac Objects

J/AJ/104/92/  fig34567.dat table0.dat  (48Kb)

Scowen P.A., Dufour R.J., Hester J.J. Astron. J. 104, 92 (1992) The HII Regions of the Galaxy M101

J/AJ/104/340 See II/183A/

J/AJ/104/704/  table1.dat table2.dat table3.dat  (33Kb)

Sramek R.A., Cowen J.J., Roberts D.A., Goss W.M., Ekers R.D. Astron. J. 104, 704 (1992) A VLA Search for Young Galactic Supernova Remnants

J/AJ/104/862/  table2  (17Kb)

McDonald S.W., Elliot J.L. Astron. J. 104, 862 (1992) Triton stellar occultation candidates 1992-1994

J/AJ/104/1349/  table1  (8Kb)

FREEDMAN W.L. Astron. J. 104, 1349 (1992) Infrared-luminous giants in M32: an intermediate-age population?

J/AJ/104/1706/  notes2 notes3 table2 table3  (33Kb)

CRAMPTON D., COWLEY A.P., HARTWICK F.D.A., KO P.W. Astron. J. 104, 1706 (1992) Canada-France-Hawaii Telescope blue grens quasar candidates

J/AJ/104/1765/  table4.dat table5.dat  (95Kb)

von Hippel T.A. Astron. J. 104, 1765 (1992) Galactic structure from faint Stroemgren photometry: the catalog of observations

J/AJ/104/2078/  table1  (7Kb)

OEGERLE W.R., HILL J.M. Astron. J. 104, 2078 (1992) (From AAS CD-ROM series, Volume 1, 1993) Structure, rotation, and the peculiar velocity cD galaxy in Abell 2107

J/AJ/105/35/  table1  (13Kb)

COZIOL R., DEMERS S., PENA M., TORRES-PEIMBERT S., FONTAINE G., WESEMAEL F., LAMON-TAGNE R. Astron. J. 105, 35 (1993) (From AAS CD-ROM series, Volume 1, 1993) The Montreal Blue Galaxy Survey. I. First list of ultraviolet-bright candidates

J/AJ/105/226/  table1 table3 table5 table7 table10 table11  (36Kb)

SODERBLOM D.R., MAYOR M. Astron. J. 105, 226 (1993) (From AAS CD-ROM series, Volume 1, 1993) Stellar kinematic groups. I. The Ursa Major Group

J/AJ/105/258/  table1 table2 table4 table5 table6 tables.tex  (78Kb)

MARRACO H.G., VEGA E.I., VRBA F.J. Astron. J. 105, 258 (1993) (From AAS CD-ROM series, Volume 1, 1993) Intracluster dust polarization in the Carina nebula

J/AJ/105/393/  table.dat  (50Kb)

Kent S.M., Ramella M., Nonino M. Astron. J. 105, 393 (1993) A Small Drift Scan Survey for Galaxies in the Northern Sky

J/AJ/105/427/  table1.dat table2.dat  (22Kb)

Rego M., Cordero-Gracia M., Zamorano J., Gallego J. Astron. J. 105, 427 (1993) IRAS observations of H-alpha selected emission-line galaxies.

J/AJ/105/499/  table1.dat  (171Kb)

REGAN M.W., WILSON C.D. Astron. J. 105, 499 (1993) OB Associations and Spiral Structure in the Southern Arm of M33

J/AJ/105/527/  table3 table4  (30Kb)

Walker A.R. Astron. J. 105, 527 (1993) (From AAS CD-ROM series, Volume 1, 1993) The Large Magellanic Cloud cluster NGC 1835: photometry of the RR Lyrae stars

J/AJ/105/788/  table1  (22Kb)

Fabricant D., Kurtz M., Geller M., Zabludoff A., Mack P., Wegner G. Astron. J. 105, 788 (1993) A study of the rich cluster of galaxies A119.

J/AJ/105/853/  table12  (17Kb)

IMPEY C., GREGORINI L. Astron. J. 105, 853 (1993) Energy distributions of radio galaxies

J/AJ/105/894/  table2a.dat table2b.dat table3a.dat table3b.dat table3c.dat  (180Kb)

GREGGIO L., MARCONI G., TOSI M., FOCARDI P. Astron. J. 105, 894 (1993) (From AAS CD-ROM series, Volume 1, 1993) Star formation in dwarf irregular galaxies: DDO 210 and NGC 3109

J/AJ/105/938/  table1  (25Kb)

FISCHER P., WELCH D.L., MATEO M. Astron. J. 105, 938 (1993) Dynamics of the young binary LMC cluster NGC 1850

J/AJ/105/980/  catalog.dat table4.dat table5.dat table6.dat  (60Kb)

Massey P., Johnson J. Astron. J. 105, 980 (1993) Massive stars near Eta Carinae: the stellar content of Tr 14 and Tr 16.

J/AJ/105/1196/  table1.dat table2.dat  (57Kb)

Harris H.C., Hunter D.A., Baum W.A., Jones J.H.

Astron. J. 105, 1196 (1993) (From AAS CD-ROM series, Volume 1, 1993) Photometric calibration of the HST Wide-Field/Planetary Camera. II. Ground-based observations of calibration fields

J/AJ/105/1251/  notes2.dat notes2.tex table1.dat table1.tex table2.dat table2.tex  (182Kb)

Wegner G., Haynes M.P., Giovanelli R. Astron. J. 105, 1251 (1993) (From AAS CD-ROM series, Volume 1, 1993) A survey of the Pisces-Perseus supercluster. V. The declination strip +33.5 deg to +39.5 deg and the main supercluster ridge

J/AJ/105/1271/  notes1.dat table1.dat table1.tex  (284Kb)

Giovanelli R., Haynes M.P. Astron. J. 105, 1271 (1993) A survey of the Pisces-Perseus supercluster. VI. The declination zone +15.5 deg to +21.5 deg

J/AJ/105/1441/  table1 table2  (29Kb)

PROSSER C.F. Astron. J. 105, 1441 (1993) (From AAS CD-ROM series, Volume 1, 1993) The open cluster IC 4665

J/AJ/105/1571/  notes1.dat table1 table2  (43Kb)

Harrington R.S., Dahn C.C., Kallarakal V.V., Guetter H.H., Riepe B.Y., Walker R.L., Pier J.R., Vrba F.J., Luginbuhl C.B., Harris H.C., Ables H.D. Astron. J. 105, 1571 (1993) (From AAS CD-ROM series, Volume 1, 1993) U. S. Naval Observatory photographic parallaxes. List IX

J/AJ/105/1637/  notes1.txt table1.dat table2.dat tables.tex  (172Kb)

Huchra J., Latham D.W., Nicolaci da Costa L., Pellegrini P.S., Willmer C.N.A. Astron. J. 105, 1637 (1993) (From AAS CD-ROM series, Volume 1, 1993) The morphological catalogue of galaxies equatorial survey

J/AJ/105/1666 See VIII/27/

J/AJ/105/1762/  table3  (32Kb)

OSTROV P., GEISLER D., FORTE J.C. Astron. J. 105, 1762 (1993) (From AAS CD-ROM series, Volume 1, 1993) The metallicity gradient and distribution function of globular clusters around NGC 1399

J/AJ/105/1822/  table3.dat  (42Kb)

Cudworth K.M., Martin S.C., Degioia-Eastwood K. Astron. J. 105, 1822 (1993) (From AAS CD-ROM series, Volume 1, 1993) Proper motions, membership, and photometry of open clusters near eta Carinae

J/AJ/105/1860/  table1  (28Kb)

COHEN M. Astron. J. 105, 1860 (1993) (From AAS CD-ROM series, Volume 1, 1993) A model of the 2-35 micron point source infrared sky

J/AJ/105/1962/  notes.dat table1.dat  (62Kb)

Weis E.W. Astron. J. 105, 1962 (1993) (From AAS CD-ROM series, Volume 1, 1993) Photometry of dwarf K and M stars

J/AJ/105/2265/  table1  (7Kb)

FEKEL F.C., BROWNING J.C., HENRY G.W., MORTON M.D., HALL D.S. Astron. J. 105, 2265 (1993) (From AAS CD-ROM series, Volume 1, 1993) Chromospherically active stars. X. Spectroscopy and photometry of HD 212280

J/AJ/105/2291/  table1 table1.tex  (36Kb)

TRUDEL J.-L., FERNIE J.D., MOCHNACKI S. Astron. J. 105, 2291 (1993) (From AAS CD-ROM series, Volume 1, 1993) The eclipsing binary EG Serpentis

J/AJ/105/2299/  table1 table2  (14Kb)

SODERBLOM D.R., PILACHOWSKI C.A., FEDELE S.B., JONES B.F. Astron. J. 105, 2299 (1993) (From AAS CD-ROM series, Volume 1, 1993) The evolution of the lithium abundances of solar-type stars. II. The Ursa Major group

J/AJ/105/2353 See I/202/

J/AJ/106/56/  table  (41Kb)

KNAPEN J.H., ARNTH-JENSEN N., CEPA J., BECKMAN J.E. Astron. J. 106, 56 (1993) (From AAS CD-ROM series, Volume 1, 1993)

Statistics and properties of HII regions in NGC 6814

J/AJ/106/66/  table2  (32Kb)

Lee M.G., Madore B.F Astron. J. 106, 66 (1993) BVRI CCD photometry of the resolved galaxy NGC 1560

J/AJ/106/154/  table1 table2 table3  (21Kb)

Ferraro F.R., Paresce F. Astron. J. 106, 154 (1993) Hubble Space Telescope / Faint Object Camera observations of the inner regions of the galactic globular cluster M15.

J/AJ/106/181/  table3.dat table4.dat table5.dat  (214Kb)

Montgomery K.A., Marschall L.A., Janes K.A. Astron. J. 106, 181 (1993) CCD photometry of the old open cluster M67

J/AJ/106/220/  table1  (35Kb)

GUETTER H.H. Astron. J. 106, 220 (1993) (From AAS CD-ROM series, Volume 1, 1993) Photometric study of the old galactic cluster Be 31

J/AJ/106/352/  notes1 notes1.tex table1  (96Kb)

Hartkopf W.I., Mason B.D., Barry D.J., McAlister H.A., Bagnuolo W.G. Astron. J. 106, 352 (1993) (From AAS CD-ROM series, Volume 1, 1993) ICCD speckle observations of binary stars. VIII. Measurements during 1989-1991 from the Cerro Tololo 4m telescope

J/AJ/106/426/  notes table1 table2 table3 table5 table6 tables.tex  (69Kb)

TYTLER D., FAN X.-M., JUNKKARINEN V.T., COHEN R.D. Astron. J. 106, 426 (1993) (From AAS CD-ROM series, Volume 1, 1993) Lick slit spectra of thirty-eight objective prism quasar candidates and low metallicity halo stars

J/AJ/106/493/  table3  (15Kb)

LEE M.G., GEISLER D. Astron. J. 106, 493 (1993) (From AAS CD-ROM series, Volume 1, 1993) Metal abundances for a large sample of globular clusters in M87

J/AJ/106/530/  format.tex table1 table2 table3 table4 table5 tables.tex  (53Kb)

BOTHUN G.D., SCHOMBERT J.M., IMPEY C.D., SPRAYBERRY D., MCGAUGH S.S. Astron. J. 106, 530 (1993) (From AAS CD-ROM series, Volume 1, 1993) The small scale environment of low surface brightness disk galaxies

J/AJ/106/560 See II/187/

J/AJ/106/591/  table2  (27Kb)

RODGERS A.W., ROBERTS W.H., WALKER I. Astron. J. 106, 591 (1993) (From AAS CD-ROM series, Volume 1, 1993) Early-type stars in orthogonal halo fields. I. An objective-prism survey

J/AJ/106/604/  table2  (14Kb)

HARRIS H.C. Astron. J. 106, 604 (1993) The globular cluster NGC 6366: its blue stragglers and variable stars

J/AJ/106/637/  table3  (9Kb)

Mason B. D., Hartkopf W.I., McAlister H.A., Sowell J.R. Astron. J. 106, 637 (1993) (From AAS CD-ROM series, Volume 1, 1993) ICCD speckle observations of binary stars. IX. A duplicity survey of the Pleiades, Praesepe, and IC 4665 clusters

J/AJ/106/656/  table2 table2a.tex table2b.tex  (25Kb)

BROOKE T.Y., TOKUNAGA A.T., STROM S.E. Astron. J. 106, 656 (1993) (From AAS CD-ROM series, Volume 1, 1993) Dust emission features in 3 micron spectra of Herbig Ae/Be stars

J/AJ/106/773/  table1 table1.tex table2 table2.tex table5 table5.tex  (51Kb)

Henry T.J., McCarthy D.W. Jr. Astron. J. 106, 773 (1993) (From AAS CD-ROM series, Volume 1, 1993) The mass-luminosity relation for stars of mass 1.0 to 0.08 solar masses

J/AJ/106/831/  tables tables.tex  (79Kb)

HILL J.M., OEGERLE W.R., Astron. J. 106, 831 (1993) (From AAS CD-ROM series, Volume 1, 1993) Dynamics of cD clusters of galaxies. I. Redshift data for seven clusters

J/AJ/106/848/  table1 table3a table3b  (41Kb)

Drinkwater M.J., Webster R.L., Thomas P.A. Astron. J. 106, 848 (1993) (From AAS CD-ROM series, Volume 1, 1993) On the nature of Mg II absorption line systems in quasars

J/AJ/106/986/  table1  (27Kb)

ARMANDROFF T.E., DA COSTA G.S., CALDWELL N., SEITZER P. Astron. J. 106, 986 (1993) The dwarf spheroidal companions to M31: a color-magnitude diagram for And III

J/AJ/106/1024/  appendix  (386Kb)

BERGBUSCH P.A. Astron. J. 106, 1024 (1993) Luminosity functions for post-turnoff stars in globular clusters. I. NGC 288

J/AJ/106/1059/  table1 table2 table3  (31Kb)

Soderblom D.R., Jones B.F., Balachandran S., Stauffer J.R., Duncan D.K., Fedele S.B., Hudon J.D. Astron. J. 106, 1059 (1993) (From AAS CD-ROM series, Volume 1, 1993) The evolution of the lithium abundances of solar-type stars. III. The Pleiades

J/AJ/106/1080/  table2  (8Kb)

SODERBLOM D.R., FEDELE S.B., JONES B.F., STAUFFER J.R., PROSSER C.F. Astron. J. 106, 1080 (1993) (From AAS CD-ROM series, Volume 1, 1993) The evolution of the lithium abundances of solar-type stars. IV. Praesepe

J/AJ/106/1273/  table1.dat table2.dat  (47Kb)

Zabludoff A.I., Geller M.J., Huchra J.P., Vogeley M.S. Astron. J. 106, 1273 (1993) The kinematics of dense clusters of galaxies. I. The data.

J/AJ/106/1471 See II/187/

J/AJ/106/1508/  table2.dat  (37Kb)

FISCHER P., WELCH D.L., MATEO M., COTE P. Astron. J. 106, 1508 (1993) Dynamics of the globular cluster NGC 362

J/AJ/106/1524/ (5Kb) REES Jr. R.F. Astron. J. 106, 1524 (1993) (From AAS CD-ROM series, Volume 1, 1993) New proper motions in the globular cluster M5

J/AJ/106/1533/ macros.tex table1 table1a.tex table1b.tex table2 table2.tex table4 table4.tex (206Kb) HARRIS G.L.H., FITZGERALD P.M., MEHTA S., REED B.C. Astron. J. 106, 1533 (1993) (From AAS CD-ROM series, Volume 1, 1993) NGC 2287: an important intermediate-age open cluster

J/AJ/106/1547/  table2  (72Kb)

APARICIO A., ALFARO E.J., DELGADO A.J., RODRIGUEZ-ULLOA J.A., CABRERA-CANO J. Astron. J. 106, 1547 (1993) The intermediate age open cluster NGC 7044.

J/AJ/106/1639/ table1.dat table2.dat table3.dat table4.dat table5.dat table6.dat table7.dat table8.dat (62Kb) McAlister H.A., Mason B.D., Hartkopf W.I., Shara M.M., Astron. J. 106, 1639 (1993) (From AAS CD-ROM series, Volume 1, 1993) ICCD speckle observations of binary stars. X. A further survey for duplicity among the bright stars

J/AJ/106/1839/  table1.dat  (19Kb)

Rodgers A.W., Roberts W.H. Astron. J. 106, 1839 (1993) (From AAS CD-ROM series, Volume 6, 1996 Lee Brotzman [ADS] 19-Feb-1996) Early-type stars in orthogonal halo fields. II. Identification of metal-rich A stars in the halo

J/AJ/106/1906/  posmk.dat table1.dat  (76Kb)

Hillenbrand L.A., Massey P., Strom S.E., Merrill K.M. Astron. J. 106, 1906 (1993) NGC 6611: A Cluster Caught in the Act

J/AJ/106/2058/  table  (46Kb)

Chambliss C.R., Walker R.L., Karle J.H., Snodgrass H.B., Vracko Y.A. Astron. J. 106, 2058 (1993) (From AAS CD-ROM series, Volume 1, 1993) A photometric investigation of the eclipsing binary V505 Sagittarii.

J/AJ/106/2096/  table1 table2 table3a table3b table3c  (37.1 Mb)

VAN HAMME W. Astron. J. 106, 2096 (1993) (From AAS CD-ROM series, Volume 1, 1993) New limb-darkening coefficients for modeling binary star light curves

J/AJ/106/2291/  table1  (13Kb)

Reed B.C. Astron. J. 106, 2291 (1993) (From AAS CD-ROM series, Volume 1, 1993) UBV photometry of luminous stars toward the Galactic Center

J/AJ/106/2429/  table  (112Kb)

SCHMIDT E.G., REISWIG D.E. Astron. J. 106, 2429 (1993) (From AAS CD-ROM series, Volume 2, 1994) The Behlen Observatory variable star survey: finding charts and light curves for the first ninety-three stars

J/AJ/106/2493/  table1 table2  (38Kb)

KIM C., MCNAMARA D.H., CHRISTENSEN C.G. Astron. J. 106, 2493 (1993) (From AAS CD-ROM series, Volume 2, 1994) A photometric and spectrographic study of SX Phoenicis

J/AJ/106/2575/  table2.tex table3.tex table4.tex tables  (103Kb)

RUIZ M. T., TAKAMIYA M.Y., MINDEZ R., MAZA J., WISHNJEWSKY M. Astron. J. 106, 2575 (1993) (From AAS CD-ROM series, Volume 1, 1993) Proper motions in the southern ESO Areas 207, 439, AND 440

J/AJ/107/135/  table1 table2 table3 table4  (14Kb)

BALCELLS M., PELETIER R.F. Astron. J. 107, 135 (1994) Colors and color gradients in bulges of galaxies

J/AJ/107/230/  table6.dat  (54Kb)

ALCAINO G., LILLER W., ALVARADO F., WENDEROTH E. Astron. J. 107, 230 (1994) (From AAS CD-ROM series, Volume 2, 1994) BVRI CCD photometry of the globular cluster NGC 1904 (M79)

J/AJ/107/280 See I/270/

J/AJ/107/594/ notes.dat table1.dat table2.dat table3.dat table4.dat table5.dat table6.dat table7.dat table8.dat table9.dat table10.dat table11.dat table12.dat (89Kb) Eggen O.J. Astron. J. 107, 594 (1994) Evolved GK stars near the Sun. II. The young disk population.

J/AJ/107/679/  table1  (13Kb)

SMITH H.A., MATTHEWS J.M., LEE K.M., WILLIAMS J., SILBERMANN N.A., BOLTE M. Astron. J. 107, 679 (1994) (From AAS CD-ROM series, Volume 2, 1994) AH Cam: a metal-rich RR Lyrae star with the shortest known Blazhko period

J/AJ/107/751/  art8.sty decalign.sty document.tex supertab.sty table2.dat table2.tex  (139Kb)

Malina R.F., Marshall H.L., Antia B., Christian C.A., Dobson C.A., Finley D.S., Fruscione A., Girouard F.R., Hawkins I., Jelinsky P., Lewis J.W., McDonald J.S., McDonald K., Patterer R.J., Saba V.W., Sirk M.M., Stroozas B.A., Vallerga J.V., Vedder P.W., Wiercigroch A., Bowyer S. Astron. J. 107, 751 (1994) (From AAS CD-ROM series, Volume 2, 1994) The extreme ultraviolet explorer bright source list

J/AJ/107/1022/  table.tex table2 table3 table7 table14 table15  (60Kb)

RICHMOND M.W., TREFFERS R.R., FILIPPENKO A.V., PAIK Y., LEIBUNDGUT B., SCHULMAN E., COX C.V. Astron. J. 107, 1022 (1994) (From AAS CD-ROM series, Volume 2, 1994) UBVRI photometry of SN 1993J in M81: the first 120 days

J/AJ/107/1054/  aasman.sty table2.dat table2.tex  (162Kb)

Malumuth E.M., Heap S.R. Astron. J. 107, 1054 (1994) (From AAS CD-ROM series, Volume 2, 1994) UBV stellar photometry of the 30 Doradus region of the Large Magellanic Cloud with the Hubble Space Telescope

J/AJ/107/1245/  table5  (76Kb)

Schneider D.P., Schmidt M., Gunn J.E. Astron. J. 107, 1245 (1994) Spectroscopic CCD surveys for quasars at large redshift. III. The Palomar Transit Grism Survey catalog.

J/AJ/107/1381/  table4.dat table4.tex  (32Kb)

TELLO C. Astron. J. 107, 1381 (1994) (From AAS CD-ROM series, Volume 2, 1994) Sampling of globular clusters and the distance to the Galactic Center. I. Data description and analysis

J/AJ/107/1433/  table1.dat table1.tex  (42Kb)

KOBULNICKY H.A., MOLNAR L.A., JONES T.J. Astron. J. 107, 1433 (1994) (From AAS CD-ROM series, Volume 2, 1994) R band polarimetry of Cygnus OB2: implications for the magnetic field geometry and polarization models

J/AJ/107/1453/  aj pt.sty table3 table3.tex  (38Kb)

BENSON P.J., HERBST W., SALZER J.J., VINTON G., HANSON G.J., RATCLIFF S.J., WINKLER P.F., ELMEGREEN D.M., CHROMEY F., STROM C., BALONEK T.J., ELMEGREEN B.G. Astron. J. 107, 1453 (1994) (From AAS CD-ROM series, Volume 2, 1994) Light curves of SN 1993J from the Keck Northeast Astronomy Consortium

J/AJ/107/1556/  notes table2.dat  (28Kb)

GARRISON R.F., GRAY R.O. Astron. J. 107, 1556 (1994) (From AAS CD-ROM series, Volume 2, 1994) The late B-Type Stars: refined MK classification, confrontation with Stromgren photometry, and the effects of rotation

J/AJ/107/1577/  table1 table5a  (20Kb)

ANTHONY-TWAROG B.J., TWAROG B.A. Astron. J. 107, 1577 (1994) (From AAS CD-ROM series, Volume 2, 1994) Reddening estimation for halo red giants using uvby photometry

J/AJ/107/1629/  table1.dat table2.dat  (47Kb)

TARENGHI M., GARILLI B., MACCAGNI D. Astron. J. 107, 1629 (1994) (From AAS CD-ROM series, Volume 2, 1994) Galaxy structures in the Hercules region

J/AJ/107/1745/  table3.dat  (49Kb)

Yanny B., Guhathakurta P., Bahcall J.N., Schneider D.P. Astron. J. 107, 1745 (1994) (From AAS CD-ROM series, Volume 2, 1994) Globular cluster photometry with the Hubble Space Telescope. II. U, V, and I measurements of M15

J/AJ/107/1764/  table3.dat  (19Kb)

SILBERMANN N.A., SMITH H.A., BOLTE M., HAZEN M.L. Astron. J. 107, 1764 (1994) (From AAS CD-ROM series, Volume 2, 1994) RR Lyrae stars and color-magnitude diagram of the globular cluster NGC 6388

J/AJ/107/1814/  table2  (10Kb)

BUS S.J., WASSERMAN L.H., ELLIOT J.L. Astron. J. 107, 1814 (1994) Chiron stellar occultation candidates: 1993-1996

J/AJ/107/1868/  table1  (16Kb)

BERRINGTON R.C., HALL D.S. Astron. J. 107, 1868 (1994) (From AAS CD-ROM series, Volume 3, 1995) The cyclic period variability in SW Cygni

J/AJ/107/2067/  table1 table1.tex  (73Kb)

Gebhardt K., Pryor C., Williams T.B., Hesser J.E. Astron. J. 107, 2067 (1994) (From AAS CD-ROM series, Volume 3, 1995) Fabry-Perot measurements of the dynamics of globular cluster cores: M15 (NGC 7078)

J/AJ/107/2093/  table1  (17Kb)

Gieren W.P., Welch D.L., Mermilliod J.-C., Matthews J.M., Hertling G. Astron. J. 107, 2093 (1994) (From AAS CD-ROM series, Volume 3, 1995) Independent distance determinations to Milky Way Cepheids in open clusters and associations. I. The binary Cepheid DL Cas in NGC 129

J/AJ/107/2101/  table1.dat  (18Kb)

Saurer W., Seeberger R., Weinberger R., Ziener R. Astron. J. 107, 2101 (1994) Berkeley 93: a distant star cluster nestled in a dust cloud

J/AJ/107/2120/  table1  (17Kb)

Jones T.J., Mergen J., Odewahn S., Gehrz R.D., Gatley I., Merrill K.M., Probst R., Woodward C.E. Astron. J. 107, 2120 (1994) A near infrared survey of the OMC2 region

J/AJ/107/2240/  table6  (265Kb)

CARNEY B.W., LATHAM D.W., LAIRD J.B., AGUILAR L.A. Astron. J. 107, 2240 (1994) (From AAS CD-ROM series, Volume 3, 1995) A survey of proper motion stars. XII. An expanded sample

J/AJ/108/1/  refs.dat table1.dat table2.dat  (48Kb)

Struble M.F., Ftaclas C. Astron. J. 108, 1 (1994) The intrinsic shape of clusters of galaxies: correlations among flattening, richness, and velocity dispersion

J/AJ/108/111/  table5.dat table5.tex  (49Kb)

Statler T.S. Astron. J. 108, 111 (1994) (From AAS CD-ROM series, Volume 3, 1995) The Intrinsic Shape of NGC 3379

J/AJ/108/222/  table4.dat table4.tex table5.dat table5.tex table6.dat table6.tex  (66Kb)

Nemec J.M., Linnell Nemec A.F., Lutz T.E. Astron. J. 108, 222 (1994) (From AAS CD-ROM series, Volume 3, 1995) Period-luminosity-metallicity relations, pulsation modes, absolute magnitudes, and distances for population II variable stars

J/AJ/108/268/  table1.dat table2.dat  (19Kb)

Beers T.C., Bestman W., Wilhelm R. Astron. J. 108, 268 (1994) Emission-line candidates from the northern HK survey

J/AJ/108/405/  table1.dat  (19Kb)

Coziol R., Demers S., Pena M., Barneoud R. Astron. J. 108, 405 (1994) (From AAS CD-ROM series, Volume 3, 1995) The Montreal Blue Galaxy Survey. II. Second list of UV-bright candidates

J/AJ/108/538/  table1.dat table9.dat table10.dat  (27Kb)

Preston G.W., Beers T.C., Shectman S.A. Astron. J. 108, 538 (1994) The space density and kinematics of blue metal-poor main sequence stars near the solar circle

J/AJ/108/555/  table4.dat table12.dat  (204Kb)

Walker A.R. Astron. J. 108, 555 (1994) (From AAS CD-ROM series, Volume 3, 1995) BVI CCD photometry of galactic globular clusters. II. M68

J/AJ/108/585/  table.dat table3.dat  (583Kb)

Montgomery K.A., Janes K.A., Phelps R.L. Astron. J., 108, 585-593 (1994) The reddening and metallicity of NGC 6791.

J/AJ/108/896/ notes.dat table1.dat table7.dat table8.dat table10.dat table11.dat table12.dat (211Kb) MAGRI C. Astron. J. 108, 896 (1994) (From AAS CD-ROM series, Volume 3, 1995) Arm structure in normal spiral galaxies. I. Multivariate data for 492 galaxies

J/AJ/108/932/  tables  (164Kb)

Sebo K.M., Wood P.R. Astron. J. 108, 932 (1994) (From AAS CD-ROM series, Volume 3, 1995) Variable stars in and around the SMC cluster NGC 330

J/AJ/108/1016/  table9.dat  (32Kb)

Layden A.C. Astron. J. 108, 1016 (1994) (From AAS CD-ROM series, Volume 3, 1995) The metallicities and kinematics of RR Lyrae variables. I. New observations of local stars

J/AJ/108/1163/  notes.dat table1.dat  (25Kb)

Kellermann K.I., Sramek R.A., Schmidt M., Green R.F., Shaffer D.B. Astron. J. 108, 1163 (1994) The radio structure of radio loud and radio quiet quasars in the Palomar bright quasar survey.

J/AJ/108/1256/  posmk.dat table3.dat table3.tex  (107Kb)

Garmany C.D., Massey P., Parker J.W. Astron. J. 108, 1256 (1994) (From AAS CD-ROM series, Volume 3, 1995) The OB association LH 58 in the Large Magellanic Cloud

J/AJ/108/1276/  table3  (23Kb)

Mayya Y.D. Astron. J. 108, 1276 (1994) (From AAS CD-ROM series, Volume 3, 1995) Embedded clusters in giant extragalactic HII Regions. I. BVRHalpha photometry

J/AJ/108/1350/  table3.dat  (15Kb)

Whitney J.H., O’Connell R.W., Rood R.T., Dorman B., Landsman W.B., Cheng K.-P., Bohlin R.C., Hintzen P.M.N., Roberts M.S., Smith A.M., Smith E.P., Stecher T.P. Astron. J. 108, 1350 (1994) Far-ultraviolet photometry of the globular cluster Omega Cen.

J/AJ/108/1598/  tables  (29Kb)

Ferrarese L., van den Bosch F.C., Ford H.C., Jaffe W., O’Connell R.W. Astron. J. 108, 1598 (1994) (From AAS CD-ROM series, Volume 3, 1995) Hubble Space Telescope photometry of the central regions of Virgo cluster elliptical galaxies. III. Brightness profiles

J/AJ/108/1722/ table1.dat table2.dat table3.dat table4.dat table5.dat table6.dat table7.dat table8.dat tablea16.dat tablea17.dat tablea18.dat tablea19.dat tablea21.dat tablea30.dat tablea31.dat tablea32.dat tablea33.dat tablea34.dat (113Kb) Kinman T.D., Suntzeff N.B., Kraft R.P. Astron. J. 108, 1722 (1994) The structure of the galactic halo outside the solar circle as traced by the blue horizontal branch stars.

J/AJ/108/1786/  table1.dat  (319Kb)

GUHATHAKURTA P., YANNY B., BAHCALL J.N., SCHNEIDER D.P. Astron. J. 108, 1786 (1994) (From AAS CD-ROM series, Volume 3, 1995) Globular cluster photometry with the Hubble Space Telescope. III. Blue stragglers and variable stars in the core of M3

J/AJ/108/1987/  notes1.dat table1.dat table1.tex table2.dat table2.tex useful.tex  (90Kb)

Alonso M.V., da Costa L.N., Latham D.W., Pellegrini P.S., Milone A.A.E. Astron. J. 108, 1987 (1994) (From AAS CD-ROM series, Volume 4, 1995 Lee Brotzman [ADS] 05-May-1995) CCD Calibration of the Magnitude Scale for the SSRS2 Sample: The Equatorial Region

J/AJ/108/2087/ table3.dat table4.dat table5.dat table6.dat table7.dat table8.dat table9.dat table10.dat table11.dat table12.dat (45Kb) Ajhar E.A., Blakeslee J.P., Tonry J.L. Astron. J. 108, 2087 (1994) (From AAS CD-ROM series, Volume 4, 1995 Lee Brotzman [ADS] 24-Apr-1995) VRI Photometry of Globular Clusters in Virgo and Leo Ellipticals

J/AJ/108/2128/  table2.dat table3.dat table4.dat table5.dat  (41Kb)

Corwin H.G.Jr, Buta R.J., de Vaucouleurs G. Astron. J. 108, 2128 (1994) Corrections and additions to the Third Reference Catalogue of Bright Galaxies .

J/AJ/108/2299/  table2  (98Kb)

Hartkopf W.I., McAlister H.A., Mason B.D., Barry D.J., Turner N.H., Fu H.-H. Astron. J. 108, 2299 (1994) ICCD speckle observations of binary stars. XI. Measurements during 1991-1993 from the Kitt Peak 4 m telescope.

J/AJ/108/2338/  appena.dat appenb.dat notes.dat table1.dat  (100Kb)

Heintz W.D. Astron. J. 108, 2338 (1994) Photographic astrometry of binary and proper-motion stars. VIII.

J/AJ/109/14/  table2.dat table3.dat  (35Kb)

Owen F.N., Ledlow M.J., Keel W.C. Astron. J. 109, 14 (1995) Optical spectroscopy of radio galaxies in Abell clusters. I. Redshifts and emission-line properties.

J/AJ/109/173/  table3  (64Kb)

Marconi G., Tosi M., Greggio L., Focardi P. Astron. J. 109, 173 (1995) Star formation in dwarf irregular galaxies: NGC 6822.

J/AJ/109/218/  tables tables.hdr  (920Kb)

Trager S.C., King I.R., Djorgovski S. Astron. J. 109, 218 (1995) (From AAS CD-ROM series, Volume 4, 1995 Lee Brotzman [ADS] 26-Apr-1995) Catalogue of Galactic globular-cluster surface-brightness profiles

J/AJ/109/269/  table2.dat  (44Kb)

Sarajedini A., Milone A.A.E. Astron. J. 109, 269 (1995) (From AAS CD-ROM series, Volume 4, 1995 Lee Brotzman [ADS] 29-Apr-1995) BVI CCD Photometry of NGC 5053: The Most Metal-Poor Galactic Globular Cluster

J/AJ/109/359/  table3  (74Kb)

Milone E.F., Stagg C.R., Sugars B.A., McVean J.R., Schiller S.J., Kallrath D. H. Bradstreet J. Astron. J. 109, 359 (1995) (From AAS CD-ROM series, Volume 4, 1995 Lee Brotzman [ADS] 29-Apr-1995) Observations and Analysis of the Contact Binary H235 in the Open Cluster NGC 752

J/AJ/109/517/  table6  (33Kb)

Buta R., Corwin H.G.Jr., de Vaucouleurs G., de Vaucouleurs A., Longo G. Astron. J. 109, 517 (1995) Integrated photoelectric magnitudes and color indices of bright galaxies in the Johnson UBV system.

J/AJ/109/543/  table5.dat  (45Kb)

Buta R., Williams K.L. Astron. J. 109, 543 (1995) (From AAS CD-ROM series, Volume 4, 1995 Lee Brotzman [ADS] 29-Apr-1995) Total and Effective Colors of 501 Galaxies in the Cousins VRI Photometric System

J/AJ/109/618/ table1 table1.tex table2 table2.tex table3 table3.tex table4 table4.tex table5 table5.tex table6 table6.tex table7 table7.tex table8 table8.tex table9 table9.tex table10 table10.tex table11.dat table11.tex (111Kb) Nemec J.M., Mateo M., Schombert J.M. Astron. J. 109, 618 (1995) (From AAS CD-ROM series, Volume 4, 1995 Lee Brotzman [ADS] 02-May-1995) B,V Photometry of the RR Lyrae Stars in the Very Metal-Poor Globular Cluster NGC 5053

J/AJ/109/639/  table1.dat table2.dat  (53Kb)

Sosin C., King I.R. Astron. J. 109, 639 (1995) (From AAS CD-ROM series, Volume 4, 1995 Lee Brotzman [ADS] 25-Apr-1995) HST Observations of the Core of the Globular Cluster NGC 6624

J/AJ/109/650/  table2.dat table2.tex  (135Kb)

Buonanno R., Corsi C.E., Fusi Pecci F., Richer H.B., Fahlman G.G. Astron. J. 109, 650 (1995) (From AAS CD-ROM series, Volume 4, 1995 Lee Brotzman [ADS] 26-Apr-1995) Young Globular Clusters in the Milky Way: Arp 2

J/AJ/109/663/  table2.dat  (55Kb)

Buonanno R., Corsi C.E., Pulone L., Fusi Pecci F., Richer H.B., Fahlman G.C. Astron. J. 109, 663 (1995) (From AAS CD-ROM series, Volume 5, 1995 Lee Brotzman [ADS] 26-Sep-1995) Terzan 7: a young metal-rich globular cluster in the Milky Way

J/AJ/109/709/  table2.dat  (147Kb)

Ali B., Depoy D.L. Astron. J. 109, 709 (1995) A 2.2um imaging survey of the Orion A molecular cloud

J/AJ/109/817/ table.tex table1.dat table2 table2.tex table3 table3.tex table4 table4.tex table5 table5.tex (98Kb) Antokhin I., Bertrand J.-F., Lamontagne R., Moffat A.F.J., Matthews J. Astron. J. 109, 817 (1995) (From AAS CD-ROM series, Volume 4, 1995 Lee Brotzman [ADS] 26-Apr-1995) The Enigmatic WN8 Stars: Intensive Photometry of Four Southern Stars on Time Scales From 30 min to 3 Months

J/AJ/109/853/  aj pt.sty table1a.dat table1a.tex table1b.dat table1b.tex  (157Kb)

Ledlow M.J., Owen F.N. Astron. J. 109, 853 (1995) (From AAS CD-ROM series, Volume 4, 1995 Lee Brotzman [ADS] 02-May-1995) A 20 cm VLA survey of Abell clusters of galaxies. IV. The radio sample and cluster properties

J/AJ/109/874/  table2.dat  (85Kb)

Beers T.C., Kriessler J.R., Bird C.M., Huchra J.P. Astron. J. 109, 874 (1995) (From AAS CD-ROM series, Volume 4, 1995 Lee Brotzman [ADS] 03-May-1995) Kinematics and Dynamics of the MKW/AWM Poor Clusters

J/AJ/109/960/  table1.dat  (51Kb)

Whitmore B.C., Schweizer F.

Astron. J. 109, 960 (1995) (From AAS CD-ROM series, Volume 4, 1995 Lee Brotzman [ADS] 03-May-1995) Hubble Space Telescope observations of young star clusters in NGC 4038/4039, ”The Antennae” galaxies

J/AJ/109/1055/  table6.dat  (281Kb)

Perelmuter J.M., Racine M. Astron. J. 109, 1055 (1995) (From AAS CD-ROM series, Volume 4, 1995 Lee Brotzman [ADS] 03-May-1995) The Globular Cluster System of M81

J/AJ/109/1086/  table1.dat table2.dat  (308Kb)

Sarajedini A., Layden A.C. Astron. J. 109, 1086 (1995) (From AAS CD-ROM series, Volume 5, 1995 Lee Brotzman [ADS] 17-Jul-1995) A Photometric Study of the Globular Cluster M54 and the Sagittarius Dwarf Galaxy: Evidence for Three Distinct Populations

J/AJ/109/1095/  table2.dat table4.dat  (17Kb)

Gilmore G., Wyse R.F.G., Jones J.B. Astron. J. 109, 1095 (1995) A determination of the thick disk chemical abundance distribution: implications for galaxy evolution.

J/AJ/109/1112/  table2.dat  (186Kb)

Sarajedini A., Forrester W.L. Astron. J. 109, 1112 (1995) (From AAS CD-ROM series, Volume 5, 1995 Lee Brotzman [ADS] 17-Jul-1995) CCD Photometry for the Galactic Globular Cluster NGC 6584

J/AJ/109/1131/  table3.dat table4.dat table5.dat table6.dat table7.dat  (133Kb)

Kuchinski L.E., Frogel J.A., Terndrup D.M., Persson S.E. Astron. J. 109, 1131 (1995) (From AAS CD-ROM series, Volume 4, 1995 Lee Brotzman [ADS] 03-May-1995) Infrared Array Photometry of Metal-Rich Globular Clusters. I. Techniques and First Results

J/AJ/109/1154/  table3.dat  (61Kb)

Frogel J.A., Kuchinski L.E., Tiede G.P. Astron. J. 109, 1154 (1995) (From AAS CD-ROM series, Volume 4, 1995 Lee Brotzman [ADS] 03-May-1995) Infrared Array Photometry of Metal Rich Globular Clusters. II. Liller 1 – The Most Metal Rich Cluster?

J/AJ/109/1269/  table.dat  (41Kb)

Etzel P.B., Olson E.C., Senay M.C. Astron. J. 109, 1269 (1995) (From AAS CD-ROM series, Volume 4, 1995 Lee Brotzman [ADS] 03-May-1995) O I Lambda 7774 Absorption as a Tracer of Streams and Disks in Nondegenerate Binary Stars

J/AJ/109/1308/  table.dat  (28Kb)

Olson E.C., Etzel P.B. Astron. J. 109, 1308 (1995) (From AAS CD-ROM series, Volume 4, 1995 Lee Brotzman [ADS] 04-May-1995) Rapid H-Alpha Emission Variations in Accretion Disks in Long-Period Algols

J/AJ/109/1352/  table2  (12Kb)

McDonald S.W., Elliot J.L. Astron. J. 109, 1352 (1995) Triton stellar occultation candidates: 1995-1999

J/AJ/109/1363/  table  (27Kb)

Wasserman L.H., Millis R.L., Bowell E. Astron. J. 109, 1363 (1995) Occultations of stars by asteroids in 1995 and 1996.

J/AJ/109/1379/  table1 table2 table3 table4  (30Kb)

Allen L.E., Strom K.M. Astron. J. 109, 1379 (1995) Moderate-resolution spectral standards from 560 to 900 nm

J/AJ/109/1458/  table2 table2.tex table3 table3.tex  (76Kb)

Ramella M., Geller M.J., Huchra J.P., Thorstensen J.R. Astron. J. 109, 1458 (1995) The redshift-space neighbourhoods of 36 loose groups of galaxies. I. The data.

J/AJ/109/1498/  table4.dat  (85Kb)

Hewett P.C., Foltz C.B., Chaffee F.H. Astron. J. 109, 1498 (1995) The Large Bright Quasar Survey. VI. Quasar catalog and survey parameters.

J/AJ/109/1608/ ngc4192.dat ngc4216.dat ngc4254.dat ngc4303.dat ngc4321.dat ngc4394.dat ngc4438.dat ngc4457.dat ngc4501.dat ngc4535.dat ngc4548.dat ngc4567.dat ngc4568.dat ngc4639.dat ngc4651.dat (27Kb) Rauscher B.J.

Astron. J. 109, 1608 (1995) (From AAS CD-ROM series, Volume 4, 1995 Lee Brotzman [ADS] 04-May-1995) Near-Infrared Surface Photometry and Morphology in Virgo Cluster Spiral Galaxy Nuclear Regions

J/AJ/109/1797/  table1.dat table1.tex  (61Kb)

Eaton J.A. Astron. J. 109, 1797 (1995) (From AAS CD-ROM series, Volume 4, 1995 Lee Brotzman [ADS] 04-May-1995) H-Alpha Measurements for Cool Giants

J/AJ/109/1903/  table1  (16Kb)

Missana M., Missana N. Astron. J. 109, 1903 (1995) Amendment of our method for deriving stellar membership of an open cluster

J/AJ/109/2318/  table1.dat table2.dat  (75Kb)

Condon J.J., Anderson E., Broderick J.J. Astron. J. 109, 2318 (1995) (From AAS CD-ROM series, Volume 5, 1995 Lee Brotzman [ADS] 17-Jul-1995) Radio identifications of extragalactic IRAS sources

J/AJ/109/2368/  table1 table3 tables.tex  (39Kb)

Thorstensen J.R., Kurtz M.J., Geller M.J., Ringwald F.A., Wegner G. Astron. J. 109, 2368 (1995) Redshifts for fainter galaxies in the First CfA Slice. III. To the Zwicky catalog limit.

J/AJ/109/2376/  table2.dat  (20Kb)

Salzer J.J., Moody J.W., Rosenberg J.L., Gregory S.A., Newberry M.V. Astron. J. 109, 2376 (1995) Imaging and spectroscopic observations of the Case Survey blue/emission-line galaxies.

J/AJ/109/2480/ field1.fit field2.fit field3.fit field4.fit field5.fit table2.dat table2.tex table7.dat table7.tex table8.dat table8.tex (40.8 Mb) Brewer J.P., Richer H.B., Crabtree D.R. Astron. J. 109, 2480 (1995) (From AAS CD-ROM series, Volume 5, 1995 Lee Brotzman [ADS] 26-Sep-1995) Late-type stars in M31. I. Photometric study of AGB stars and metallicity gradients

J/AJ/109/2503/ table3a.dat table3b.dat table3c.dat table4a.dat table4b.dat table4c.dat table5a.dat table5b.dat table5c.dat (58Kb) Mayya Y.D. Astron. J. 109, 2503 (1995) (From AAS CD-ROM series, Volume 5, 1995 Lee Brotzman [ADS] 19-Jul-1995) Embedded clusters in giant extragalactic H II regions. II. Evolutionary population synthesis model

J/AJ/109/2828/  catalog  (162Kb)

Twarog B.A., Anthony-Twarog B.J. Astron. J. 109, 2828 (1995) Ca II H and K filter photometry on the uvby system. II. The catalog of observations.

J/AJ/110/212/  table3  (89Kb)

Aparicio A., Cepa J., Gallart C., Castaneda H.O., Chiosi C., Bertelli G., Mas-Hesse J.M., Munoz-Tunon C., Telles E., Tenorio-Tagle G., Diaz A.I., Garcia-Vargas L.M., Garzon F., Gonzalez-Delgado R.M., Perez E., Rodriguez-Espinosa J.M., Terlevich E., Televich R.J., Varela A.M., Vilchez J.M. Astron. J. 110, 212 (1995) Bright stars and recent star formation in the irregular Magellanic galaxy NGC 2366.

J/AJ/110/268/  table1 table2 table3  (17Kb)

Creese M., Jones T.J., Kobulnicky H.A. Astron. J. 110, 268 (1995) Red and infrared polarimetry of highly reddened stars in the galactic plane.

J/AJ/110/275/ alp cma.dat alp cma.hdr alp lyr.dat alp lyr.hdr table4.dat table4.hdr table5.dat table5.hdr table6.dat table6.hdr table7.dat table7.hdr table8.dat table8.hdr (465Kb) Cohen M., Witteborn F.C., Walker R.G., Bregman J.D., Wooden D.H. Astron. J. 110, 275 (1995) (From AAS CD-ROM series, Volume 5, 1995 Lee Brotzman [ADS] 26-Oct-1995) Spectral irradiance calibration in the infrared. IV. 1.2-35 um spectra of six standard stars

J/AJ/110/463/  table2.dat  (46Kb)

Quintana H., Ramirez A., Melnick J., Raychaudhury S., Slezak E. Astron. J. 110, 463 (1995) (From AAS CD-ROM series, Volume 5, 1995 Lee Brotzman [ADS] 06-Oct-1995) The Shapley supercluster. I. Spectroscopic observations in the central region

J/AJ/110/638/  table5.dat  (488Kb)

Walker A.R. Astron. J. 110, 638 (1995) (From AAS CD-ROM series, Volume 5, 1995 Lee Brotzman [ADS] 27-Sep-1995)

The LMC cluster NGC 1866. I. A revised photometric sequence and a reconnaissance of the surrounding field population

J/AJ/110/652/  tablecd1.dat tablecd2.dat tablecd3.dat  (276Kb)

Fullton L.K., Carney B.W., Olszewski E.W., Zinn R., Demarque P., Janes K.A., Da Costa G.S., Seitzer P. Astron. J. 110, 652 (1995) (From AAS CD-ROM series, Volume 5, 1995 Lee Brotzman [ADS] 27-Sep-1995) A VIc color-magnitude diagram of the globular cluster NGC 6352 from Hubble Space Telescope Planetary Camera observations

J/AJ/110/704/ table2 table3 table4a table4b table4c table4d table5a table5b table5c table5d table6 (133Kb) Silbermann N.A., Smith Horace A. Astron. J. 110, 704 (1995) The RR Lyrae Variable Stars in the Globular Cluster M15

J/AJ/110/880/  table4.dat table5.dat table6.dat table9.dat  (155Kb)

Johnston K.J., Fey A.L., Zacharias N., Russell J.L., Ma C., de Vegt C., Reynolds J.E., Jauncey D.L., Archinal B.A., Carter M. S., Corbin T.E., Eubanks M.T., Florkowski D.R., Hall D.M., McCarthy D.D., McCulloch P.M., King E.A., Nicolson G., Shaffer D.B. Astron. J. 110, 880 (1995) (From AAS CD-ROM series, Volume 5, 1995 Lee Brotzman [ADS] 27-Sep-1995) A radio reference frame

J/AJ/110/1032/  table1  (14Kb)

Muriel H., Nicotra M.A., Lambas D.G. Astron. J. 110, 1032 (1995) The luminosity function of elliptical galaxies

J/AJ/110/1083/  table1.dat table1.tex  (180Kb)

Takamiya M., Kron R.G., Kron G.E. Astron. J. 110, 1083 (1995) (From AAS CD-ROM series, Volume 5, 1995 Lee Brotzman [ADS] 27-Sep-1995) Photoelectric photometry of Zwicky Galaxies

J/AJ/110/1141/  table5a.dat table5b.dat table5c.dat table5d.dat  (38Kb)

Mateo M., Kubiak M., Szymanski M., Kaluzny J., Krzeminski W., Udalski A. Astron. J. 110, 1141 (1995) (From AAS CD-ROM series, Volume 5, 1995 Lee Brotzman [ADS] 30-Oct-1995) The Optical Gravitational Lensing Experiment: variable stars in the Sagittarius dwarf spheroidal galaxy

J/AJ/110/1186/ table3.dat table4.dat table5.dat table6.dat table7.dat table8.dat table9.dat table10.dat table11.dat table12.dat table13.dat table14.dat table15.dat table16.dat table17.dat table18.dat tables.tex (128Kb) Nemec J.M., Mateo M., Burke M., Olszewski E.W. Astron. J. 110, 1186 (1995) (From AAS CD-ROM series, Volume 5, 1995 Lee Brotzman [ADS] 30-Oct-1995) SX Phoenecis stars in the extremely metal-poor globular cluster NGC 5053

J/AJ/110/1229/  table4 table5 table8  (33Kb)

Prosser C.F., Stauffer J.R., Caillault J.-P., Balachandran S., Stern R.A., Randich S. Astron. J. 110, 1229 (1995) An X-ray survey of the open cluster NGC 6475 (M7) with ROSAT.

J/AJ/110/1350/  table2 table3 table4 table5.dat  (109Kb)

Van Hamme W., Wilson R.E., Guinan E.F. Astron. J. 110, 1350 (1995) (From AAS CD-ROM series, Volume 5, 1995 Lee Brotzman [ADS] 28-Sep-1995) Periodic light curve changes for Beta Lyrae

J/AJ/110/1420/  table2.dat  (16Kb)

Williams J.G. Astron. J. 110, 1420 (1995) (From AAS CD-ROM series, Volume 5, 1995 Lee Brotzman [ADS] 30-Oct-1995) Planetary-induced nutation of the Earth: direct terms

J/AJ/110/1649/  table1.dat table1.fit  (458Kb)

Mighell K.J., Rich R.M. Astron. J. 110, 1649 (1995) (From AAS CD-ROM series, Volume 5, 1995 Lee Brotzman [ADS] 31-Oct-1995) Hubble Space Telescope Planetary Camera observations of the stellar populations near the nucleus of M33

J/AJ/110/1699/  table1a table1b table1c  (49Kb)

Gebhardt K., Pryor C., Williams T.B., Hesser J.E. Astron. J. 110, 1699 (1995) (From AAS CD-ROM series, Volume 5, 1995 Lee Brotzman [ADS] 31-Oct-1995) Fabry-Perot observations of globular clusters. II. 47 Tuc, NGC 6397, and M30

J/AJ/110/1774/  table2.dat table2.tex table7.dat table7.tex  (185Kb)

Terndrup D.M., Sadler E.M., Rich R.M.

Astron. J. 110, 1774 (1995) (From AAS CD-ROM series, Volume 5, 1995 Lee Brotzman [ADS] 02-Nov-1995) K Giants in Baade’s window. I. Velocity and line-strength measurements

J/AJ/110/1838 See III/198/

J/AJ/110/1959/  table2.dat table2.tex table3.dat table3.tex  (69Kb)

Ledlow M.J., Owen F.N. Astron. J. 110, 1959 (1995) (From AAS CD-ROM series, Volume 5, 1995 Lee Brotzman [ADS] 02-Nov-1995) A 20cm VLA survey of Abell clusters of galaxies. V. Optical observations and surface photometry

J/AJ/110/1993/ table1a.dat table1a.tex table1b.tex table2a.dat table2a.tex table2b.dat table2b.tex (169Kb) Sanders D.B., Egami E., Lipari S., Mirabel I.F., Soifer B.T. Astron. J. 110, 1993 (1995) The IRAS Bright Galaxy Survey - Part II: Extension to Southern Declinations (Dec =-30deg), and Low Galactic Latitudes (5 —b— =30deg)

J/AJ/110/2009 See VII/204/

J/AJ/110/2131/  table2s table2v table3s table3v table4s table4v table5s table5v  (62Kb)

Armandroff T.E., Olszewski E.W., Pryor C. Astron. J. 110, 2131 (1995) The mass-to-light ratios of the Draco and Ursa Minor dwarf spheroidal galaxies. I. Radial velocities from multifiber spectroscopy.

J/AJ/110/2166/  table2 table8 table9 table10  (284Kb)

Mateo M.,Fischer P.,Krzeminski W. Astron. J. 110, 2166 (1995) (From AAS CD-ROM series, Volume 5, 1995 Lee Brotzman [ADS] 03-Nov-1995) Variable Stars in the Sextans Dwarf Spheroidal Galaxy

J/AJ/110/2206/  b-vphot bphot image.fit table1.dat table14.dat vphot  (2.3 Mb)

Kaluzny J., Krzeminski W., Mazur B. Astron. J. 110, 2206 (1995) (From AAS CD-ROM series, Volume 6, 1996 Lee Brotzman [ADS] 19-Feb-1996) CCD Photometry of Variable Stars in the Globular Cluster Ru 106

J/AJ/110/2288/  table2.dat  (18Kb)

Layden A.C. Astron. J. 110, 2288 (1995) (From AAS CD-ROM series, Volume 5, 1995 Lee Brotzman [ADS] 03-Nov-1995) The metallicities and kinematics of RR Lyrae variables. II. Galactic structure and formation from local stars

J/AJ/110/2319/  table3.dat  (56Kb)

Clementini G., Carretta E., Gratton R., Merighi R., Mould J.R. McCarthy J.K. Astron. J. 110, 2319 (1995) (From AAS CD-ROM series, Volume 6, 1996 Lee Brotzman [ADS] 19-Feb-1996) The composition of HB stars: RR Lyrae variables

J/AJ/110/2400/  table1.dat  (32Kb)

Nelson R.H., Milone E.F., VanLeeuwen J., Terrell D., Penfold J.E., Kallrath J. Astron. J. 110, 2400 (1995) (From AAS CD-ROM series, Volume 6, 1996 Lee Brotzman [ADS] 1-May-1996) Observations and analysis of the field contact binary V728 Herculis

J/AJ/110/2715/  m33.dat ngc6822.dat  (49Kb)

Massey P., Armandroff T.E., Pyke R., Patel K., Wilson C.D. Astron. J. 110, 2715 (1995) (From AAS CD-ROM series, Volume 6, 1996 Lee Brotzman [ADS] 19-Feb-1996) Hot, luminous stars in selected regions of NGC 6822, M31 and M33

J/AJ/110/2739/  table1 table2 table4 table5 tables.tex  (69Kb)

Calzetti D., Kinney A.L., Ford H., Doggett J., Long K.S. Astron. J. 110, 2739 (1995) An atlas of H-alpha emitting regions in M33: a systematic search for SS433 star candidates.

J/AJ/110/2747/  table3.dat table4.dat  (143Kb)

Schweitzer A.E., Cudworth K.M., Majewski S.R., Suntzeff N.B. Astron. J. 110, 2747 (1995) (From AAS CD-ROM series, Volume 7, 1996 Lee Brotzman [ADS] 14-Jul-1996) The absolute proper motion and a membership survey of the Sculptor dwarf spheroidal galaxy

J/AJ/110/2788/  table3.dat table4.dat table5.dat table6.dat tables.tex  (117Kb)

Tiede G.P., Frogel J.A., Terndrup D.M. Astron. J. 110, 2788 (1995) Implications of new JHK photometry and a deep infrared luminosity function for the galactic bulge.

J/AJ/110/2844/  table2.dat table3.dat table6.dat table7.dat table8.dat  (28Kb)

Kuchinski L.E., Frogel J.A. Astron. J. 110, 2844 (1995)

Infrared array photometry of metal-rich globular clusters. III. Two more clusters and an analysis of V-K colors.

J/AJ/111/42/  table3  (14Kb)

Tomita A., Nakamura F.E., Takata T., Nakanishi K., Takeuchi T., Ohta K., Yamada T. Astron. J. 111, 42 (1996) Distribution of blue galaxies in a merging cluster of galaxies Abell 168.

J/AJ/111/370/  table1  (16Kb)

Hartkopf W.I., Mason B.D., McAlister H.A. Astron. J. 111, 370 (1996) (From AAS CD-ROM series, Volume 6, 1996 Lee Brotzman [ADS] 24-Feb-1996) Binary star orbits from speckle interferometry. VIII. Orbits of 37 close visual systems

J/AJ/111/393/  table1.dat table3  (85Kb)

Al-Shukri A.M., McAlister H.A., Hartkopf W.I., Hutter D.J., Franz O.G. Astron. J. 111, 393 (1996) (From AAS CD-ROM series, Volume 6, 1996 Lee Brotzman [ADS] 27-Feb-1996) ICCD Speckle observations of binary stars. XII. Measurements during 1984-1986 from the Perkins 1.8 m telescope

J/AJ/111/439/  notes.dat table4.dat  (127Kb)

Henry T.J., Soderblom D.R., Donahue R.A., Baliunas S.L. Astron. J. 111, 439 (1996) (From AAS CD-ROM series, Volume 6, 1996 Lee Brotzman [ADS] 07-Mar-1996) A survey of Ca II H and K chromospheric emission in southern solar-type stars

J/AJ/111/645/  notes.dat table1.dat table1.tex  (29Kb)

Fabian D., Usher P.D. Astron. J. 111, 645 (1996) (From AAS CD-ROM series, Volume 6, 1996 Lee Brotzman [ADS] 07-Mar-1996) Optical Morphology of Bright Quasars on Images Taken With 1.2 m Schmidt Cameras

J/AJ/111/696/  table2  (90Kb)

Keel W.C. Astron. J. 111, 696 (1996) (From AAS CD-ROM series, Volume 6, 1996 Lee Brotzman [ADS] 08-Mar-1996) Seyfert galaxies with companions: orbital and kinematic clues to AGN triggering

J/AJ/111/846/  table1.dat table1.ps table2.dat table2.ps  (117Kb)

O’Dell C.R., Wong S.K. Astron. J. 111, 846 (1996) Hubble Space Telescope mapping of the Orion nebula. I. A survey of stars and compact objects

J/AJ/111/902/  table3.dat table3.tex table4.dat table4.tex  (146Kb)

Henden A.A. Astron. J. 111, 902 (1996) (From AAS CD-ROM series, Volume 6, 1996 Lee Brotzman [ADS] 02-May-1996) Faint Cepheid studies. I. Comparison with existing data sets

J/AJ/111/936/  table1.dat table2.dat  (77Kb)

Hartkopf W.I., Mason B.D., McAlister H.A., Turner N.H., Barry D.J., Franz O.G., Prieto C.M. Astron. J. 111, 936 (1996) (From AAS CD-ROM series, Volume 6, 1996 Lee Brotzman [ADS] 10-Mar-1996) ICCD Speckle observations of binary stars. XIII. Measurements during 1989-1994 from the Cerro Tololo 4 m telescope

J/AJ/111/1128/  table1.dat  (41Kb)

Malumuth E.M., Waller W.H., Parker J.Wn. Astron. J. 111, 1128 (1996) (From AAS CD-ROM series, Volume 6, 1996 Lee Brotzman [ADS] 02-May-1996) The stellar initial mass function in the giant H II region NGC 595

J/AJ/111/1193/  table5.dat  (48Kb)

Jones B.F., Prosser C.F. Astron. J. 111, 1193 (1996) Membership of stars in NGC 1039 (M34)

J/AJ/111/1596/  table.tex table3.dat table4.dat  (59Kb)

Smith E.O., Neill J.D., Mighell K.J., Rich M. Astron. J. 111, 1596 (1996) (From AAS CD-ROM series, Volume 6, 1996 Lee Brotzman [ADS] 03-May-1996) Fornax globular clusters 1 and 5: A confirmed extragalactic second parameter Pair

J/AJ/111/1945 See VIII/42/

J/AJ/111/1964/  sqiid.dat  (22Kb)

Lada C.J., Alves J., Lada E.A. Astron. J., 111, 1964-1976 (1996) Near-infrared imaging of embedded clusters: NGC 1333.

J/AJ/112/36/  a3266.dat table1  (63Kb)

Quintana H., Ramirez A., Way M.J. Astron. J. 112, 36 (1996) Dynamics of the cluster of galaxies A3266 (SERSIC 40/6). I. Spectroscopic data

J/AJ/112/62/  table1 table2 table3 table4 table5 table6  (29Kb)

Tanner A.M., Bechtold J., Walker C.E., Black J.H., Cutri R.M. Astron. J. 112, 62 (1996) A study of quasar absorption-line systems with IRAS

J/AJ/112/171/  table1.dat table1.tex  (45Kb)

Sadler E.M., Rich R.M., Terndrup D.M. Astron. J. 112, 171 (1996) (From AAS CD-ROM series, Volume 7, 1996 Lee Brotzman [ADS] 09-Oct-1996) K Giants in Baade’s Window. II. The Abundance Distribution

J/AJ/112/241/  table2.dat table3.dat table4.dat  (203Kb)

Cohen M., Witteborn F.C., Bregman J.D., Wooden D.H., Salama A., Metcalfe L. Astron. J. 112, 241 (1996) (From AAS CD-ROM series, Volume 7, 1996 Lee Brotzman [ADS] 03-Oct-1996) Spectral irradiance calibration in the infrared. VI. 3-35 um spectra of three southern standard stars

J/AJ/112/359/  aj pt ha.sty table1.dat table1.tex  (100Kb)

van den Bergh S., Abraham R.G., Ellis R.S., Tanvir N.R., Santiago B.X., Glazebrook K.G. Astron. J. 112, 359 (1996) (From AAS CD-ROM series, Volume 7, 1996 Lee Brotzman [ADS] 03-Oct-1996) A morphological catalog of galaxies in the Hubble Deep Field

J/AJ/112/407/  table1.dat table2.dat table3.dat table4.dat  (20Kb)

Gregg M.D., Becker R.H., White R.L., Helfand D.J., McMahon R.G., Hook I.M. Astron. J. 112, 407 (1996) The FIRST bright QSO survey.

J/AJ/112/491/ cfht01.eps cfht02.eps cfht03.eps cfht04.eps cfht05.eps cfht06.eps cfht07.eps cfht08.eps cfht09.eps cfht10.eps cfht11.eps cfht12.eps cfht13.eps cfht14.eps cfht15.eps cfht16.eps cfht17.eps cfht18.eps cfht19.eps cfht20.eps cfht21.eps cfht22.eps cfht23.eps cfht24.eps cfht25.eps cfht26.eps cfht27.eps cfht28.eps cfht29.eps dao01.eps dao02.eps dao03.eps dao04.eps dao05.eps (4.2 Mb) Brewer J.P., Richer H.B., Crabtree D.R. Astron. J. 112, 491 (1996) (From AAS CD-ROM series, Volume 7, 1996 Lee Brotzman [ADS] 10-Aug-1996) Late-Type Stars in M31. II. C-, S-, and M-Star Spectra

J/AJ/112/545/  table1.dat table1.tex  (40Kb)

Pilachowski C.A., Sneden C., Kraft R.P., Langer G.E. Astron. J. 112, 545 (1996) (From AAS CD-ROM series, Volume 7, 1996 Lee Brotzman [ADS] 13-Aug-1996) Proton capture chains in globular cluster stars. I. Evidence for deep mixing based on sodium and magnesium abundances in M13 giants

J/AJ/112/601/  table3.dat  (17Kb)

Hill R.S., Cheng K.-P., Smith E.P., Hintzen P.M.N., Bohlin R.C., OConnell R.W., Roberts M.S., Smith A.M., Stecher T.P. Astron. J. 112, 601 (1996) (From AAS CD-ROM series, Volume 7, 1996 Lee Brotzman [ADS] 13-Aug-1996) UIT: New ultraviolet stellar photometry and surface brightness profiles of the globular cluster M 79 (NGC 1904)

J/AJ/112/628/  cat2000.dat standard.dat summary.dat tables.txt  (1.1 Mb)

Fan X., Burstein D., Chen J.S., Zhu J., Jiang Z.J., Wu H., Yan H.J., Zheng Z.Y., Zhou X., Fang L.Z., Chen F.Z., Deng Z.G., Chu Y.Q., Hester J.J., Windhorst R.A., Li Y., Lu P., Sun W-H., Chen W-P., Tsay W-S., Chiueh T-H. Astron. J. 112, 628 (1996) Deep Wide-Field Spectrophotometry of the Open Cluster M67

J/AJ/112/742/  table1 table3  (120Kb)

Wetterer C.J., McGraw J.T., Hess T.R., Grashuis R. Astron. J. 112, 742 (1996) (From AAS CD-ROM series, Volume 7, 1996 Lee Brotzman [ADS] 14-Aug-1996) RR Lyrae variable stars in the CCD/Transit Instrument Survey

J/AJ/112/772/ table1.dat table2.dat table3.dat table4.dat table5.dat table6.dat table7.dat (3.2 Mb) Caldwell J.A.R., Schechter P.L. Astron. J. 112, 772 (1996) (From AAS CD-ROM series, Volume 7, 1996 Lee Brotzman [ADS] 14-Aug-1996) Scanned CCD V and I Photometry at the South Galactic Pole

J/AJ/112/1054/  table1.dat  (95Kb)

Zinn R., Barnes S. Astron. J. 112, 1054 (1996) (From AAS CD-ROM series, Volume 7, 1996 Lee Brotzman [ADS] 25-Sep-1996) The horizontal branches of globular clusters. I. The color-magnitude diagram of NGC 6426

J/AJ/112/1061/  table6.dat  (48Kb)

Bergbusch P.A. Astron. J. 112, 1061 (1996) (From AAS CD-ROM series, Volume 7, 1996 Lee Brotzman [ADS] 25-Sep-1996) Luminosity functions for post-turnoff stars in globular clusters. II. NGC 7099

J/AJ/112/1188/  table1.dat  (19Kb)

Beers T.C., Rossi S., Ulrich D., Wilhelm R. Astron. J. 112, 1188 (1996) (From AAS CD-ROM series, Volume 7, 1996 Lee Brotzman [ADS] 09-Oct-1996) Additional emission-line candidates from the HK survey

J/AJ/112/1197/  table3.dat  (365Kb)

Brown M.E., Bouchez A.H., Spinrad H., Johns-Krull C.M. Astron. J. 112, 1197 (1996) (From AAS CD-ROM series, Volume 7, 1996 Lee Brotzman [ADS] 09-Oct-1996) A high-resolution catalog of cometary emission lines

J/AJ/112/1335/  table7.dat table7.ps tables.fit  (1.3 Mb)

Williams R.E., Blacker B., Dickinson M., van Dyke Dixon W., Ferguson H.C., Fruchter A.S., Giavalisco M., Gilliland R.L., Heyer I., Katsanis R., Levay Z., Lucas R.A., McElroy D.B., Petro L., Postman M., Adorf H.-M., Hook R.N. Astron. J. 112, 1335 (1996) The Hubble Deep Field observations, data reduction, and galaxy photometry.

J/AJ/112/1438/  table3.dat  (16Kb)

Lee M.G. Astron. J. 112, 1438 (1996) Stellar populations in the central region of the dwarf elliptical galaxy NGC 205

J/AJ/112/1487 See VII/195/

J/AJ/112/1489/  table2.tex table2a table2b  (57Kb)

Yan L., Cohen J.G. Astron. J. 112, 1489 (1996) A spectroscopic survey for binary stars in the globular cluster NGC 5053.

J/AJ/112/1803/  notes.dat refs.dat table1.dat table2.dat  (165Kb)

Marzke R.O., Huchra J.P., Geller M.J. Astron. J. 112, 1803 (1996) (From AAS CD-ROM series, Volume 8, 1997 Lee Brotzman [ADS] 30-Mar-1997) Large-Scale Structure at Low Galactic Latitude

J/AJ/112/2004/  table3  (151Kb)

Alcaino G., Liller W., Alvarado F., Kravtsov V., Ipatov A., Samus N., Smirnov O. Astron. J. 112, 2004 (1996) Multicolor CCD photometry of the SMC cluster Kron 3

J/AJ/112/2020/  table3  (59Kb)

Alcaino G., Liller W., Alvarado F., Kravtsov V., Ipatov A., Samus N., Smirnov O. Astron. J. 112, 2020 (1996) NTT photometry in UBV of the distant globular cluster NGC 1841.

J/AJ/112/2274/  table4a.dat table4b.dat table4c.dat table4d.dat table4e.dat  (192Kb)

Cohen M., Witteborn F.C., Carbon D.F., Davies J.K., Wooden D.H., Bregman J.D. Astron. J. 112, 2274 (1996) (From AAS CD-ROM series, Volume 7, 1996 Lee Brotzman [ADS] 05-Nov-1996) Spectral irradiance calibration in the infrared. VII. New composite spectra, comparison with model atmospheres, and far-infrared extrapolations

J/AJ/112/2319/  table2 table2.tex  (80Kb)

Hilton J.L., Seidelmann P.K., Middour J. Astron. J. 112, 2319 (1996) (From AAS CD-ROM series, Volume 7, 1996 Lee Brotzman [ADS] 05-Nov-1996) Prospects for determining asteroid masses

J/AJ/112/2408/  table4.dat  (37Kb)

Hamuy M., Phillips M.M., Suntzeff N.B., Schommer R.A., Maza J., Antezana R., Wischnjewsky M., Valladares G., Muena C., Gonzales L.E., Aviles R., Wells L.A., Smith R.C., Navarrete M., Covarrubias R., Williger G.M., Walker A.R., Layden A.C., Elias J.H., Baldwin J.A., Hernandez M., Tirado H., Ugarte P., Elston R., Saavedra N., Barrientos F., Costa E., Lira P., Ruiz M.T., Anguita C., Gomez X., Ortiz P., Della Valle M., Danziger J., Storm J., Kim Y-C., Bailyn C., Rubenstein E.P., Tucker D., Cersosimo S., Mendez R.A., Siciliano L., Sherry W., Chaboyer B., Koopmann R.A., Geisler D., Sarajedini A., Dey A., Tyson N., Rich R.M., Gal R., Lamontagne R., Caldwell N., Guhathakurta P., Phillips A.C., Szkody P., Prosser C., Ho L.C., McMahan R., Baggley G., Cheng K.-P., Havlen R., Wakamatsu K., Janes K., Malkan M., Baganoff F., Seitzer P., Shara M., Sturch C., Hesser J., Hartigan P., Hughes J., Welch D., Williams T.B., Ferguson H., Francis P.J., French L., Bolte M., Roth J., Odewahn S., Howell S., Krzeminski W.

Astron. J. 112, 2408 (1996) (From AAS CD-ROM series, Volume 8, 1997 Lee Brotzman [ADS] 30-Mar-1997) BVRI Light Curves for 29 Type Ia Supernovae

J/AJ/112/2471/  table2.dat table4.dat table5.dat  (60Kb)

Tully R.B., Verheijen M.A.W., Pierce M.J., Huang J.-S. , Wainscoat R.J. Astron. J. 112, 2471 (1996) (From AAS CD-ROM series, Volume 7, 1996 Lee Brotzman [ADS] 06-Nov-1996) The Ursa Major cluster of galaxies. I. Cluster definition and photometric data

J/AJ/112/2500/  table3 table4 table5 table6  (28Kb)

McLean I.S., Teplitz H. Astron. J. 112, 2500 (1996) Near infrared JHK photometry of intermediate redshift clusters

J/AJ/112/2559/  table2.dat table2.tex  (48Kb)

Sohn Y.-J., Davidge T.J. Astron. J. 112, 2559 (1996) (From AAS CD-ROM series, Volume 7, 1996 Lee Brotzman [ADS] 06-Nov-1996) Supergiant stars in the disk of NGC 672

J/AJ/112/2757/  table1 table3.dat table3.tex table4.dat table4.tex  (190Kb)

Henden A.A. Astron. J. 112, 2757 (1996) (From AAS CD-ROM series, Volume 8, 1997 Lee Brotzman [ADS] 30-Mar-1997) Faint Cepheid Studies. II. Stars with Published Periods

J/AJ/112/2769/  table1 table2  (134Kb)

Schmidt E.G., Seth A. Astron. J. 112, 2769 (1996) (From AAS CD-ROM series, Volume 7, 1996 Lee Brotzman [ADS] 06-Nov-1996) The Behlen Observatory variable star survey: Paper IV

J/AJ/112/2786/  table1 table2  (35Kb)

Richichi A., Baffa C., Calamai G., Lisi F., Astron. J. 112, 2786 (1996) The TIRGO lunar occultation program: summary of the 1985-1995 observations.

J/AJ/112/2799 See III/198/

J/AJ/112/2862 See V/98/

J/AJ/113/22/  table2 table2.tex  (214Kb)

Giovanelli R., Haynes M.P., Herter T., Vogt N.P., Wegner G., Salzer J.J., Da Costa L.N., Freudling W. Astron. J. 113, 22 (1997) (From AAS CD-ROM series, Volume 7, 1996 Lee Brotzman [ADS] 06-Nov-1996) The I Band Tully-Fisher Relation for Cluster Galaxies: Data Presentation

J/AJ/113/264/  table2.dat table3.dat table4.dat  (247Kb)

Sarajedini A., Layden A. Astron. J. 113, 264 (1997) Reddenings, Metallicities, and Possible Abundance Anomalies in Young Globular Clusters

J/AJ/113/279/  tablea.dat tablea.tex tableb.dat tableb.tex  (33Kb)

Kraft R.P., Sneden C., Smith G.H., Shetrone M.D., Langer G.E., Pilachowski C.A. Astron. J. 113, 279 (1997) (From AAS CD-ROM series, Volume 8, 1997 Lee Brotzman [ADS] 04-Apr-1997) Proton Capture Chains in Globular Cluster Stars. II. Oxygen, Sodium, Magnesium, and Aluminum Abundances in M13 Giants Brighter than the Horizontal Branch

J/AJ/113/321/  table1.dat table1.tex table3.dat table3.tex  (68Kb)

Tiede G.P., Terndrup D.M. Astron. J. 113, 321 (1997) (From AAS CD-ROM series, Volume 8, 1997 Lee Brotzman [ADS] 04-Apr-1997) A New Survey of Stellar Kinematics in the Central Milky Way

J/AJ/113/401/  table1  (57Kb)

Samec Ronald G., Pauley Brandy R., Carrigan Brian J. Astron. J. 113, 401 (1997) (From AAS CD-ROM series, Volume 8, 1997 Lee Brotzman [ADS] 04-Apr-1997) U,B,V Light Curves of the Short-Period Solar-Type Eclipsing Binary, V417 Aquilae

J/AJ/113/483/  table1.dat table1.tex  (79Kb)

Ramella M., Pisani A., Geller M.J. Astron. J. 113, 483 (1997) (From AAS CD-ROM series, Volume 8, 1997 Lee Brotzman [ADS] 04-Apr-1997) Groups of Galaxies in the Northern CfA Redshift Survey

J/AJ/113/634/  table1.tex table2a table2a.tex table2b table2b.tex table3 table3.tex  (82Kb)

Ibata R.A., Wyse R.F.G., Gilmore G., Irwin M.J., Suntzeff N.B. Astron. J. 113, 634 (1997) (From AAS CD-ROM series, Volume 8, 1997 Lee Brotzman [ADS] 04-Apr-1997) The Kinematics, Orbit, and Survival of the Sagittarius Dwarf Spheroidal Galaxy

J/AJ/113/669/  table1.dat  (275Kb)

Cohen R.L., Guhathakurta P., Yanny B., Schneider D.P., Bahcall J.N. Astron. J. 113, 669 (1997) (From AAS CD-ROM series, Volume 8, 1997 Lee Brotzman [ADS] 09-Apr-1997) Globular Cluster Photometry with the Hubble Space Telescope. VI. WF/PC-I Observations of the Stellar Populations in the Core of M13 (NGC 6205)

J/AJ/113/806/  table2.dat table4.dat  (22Kb)

Gizis J.E. Astron. J., 113, 806 (1997) M-Subdwarfs: Spectroscopic Classification and the Metallicity Scale

J/AJ/113/823 See V/92A/

J/AJ/113/905/  table1.dat table2.dat table3.dat  (61Kb)

Pantoja C.A., Altschuler D.R., Giovanardi C., Giovanelli R. Astron. J. 113, 905 (1997) 21-cm line observations of galaxies in the zone of avoidance.

J/AJ/113/950/  table2 table3 table4  (150Kb)

Fisher D. Astron. J. 113, 950 (1997) (From AAS CD-ROM series, volume 7, 1996 Lee Brotzman [ADS] 12-Nov-1996) Kinematic Profiles of S0 Galaxies

J/AJ/113/1001/  table2a.dat table2b.dat  (5.4 Mb)

Han M., Hoessel J.G., Gallagher J.S.III, Holtzman J., Stetson P.B., Trauger J., Ballester G.E., Burrows C., Clarke J., Crisp D., Griffiths R., Grillmair C., Hester J., Krist J., Mould J.R., Scowen P., Stapelfeldt K., Watson A., Westphal J. Astron. J. 113, 1001 (1997) (From AAS CD-ROM series, Volume 8, 1997 Lee Brotzman [ADS] 09-Apr-1997) Stellar Populations in the Dwarf Elliptical Galaxy NGC 147

J/AJ/113/1011/  table.dat  (659Kb)

Cornett R.H., Greason M.R., Hill J.K., Parker J.Wm., Waller W.H., Bohlin R.C., Cheng K.-P., Neff S.G., O’Connell R.W., Roberts M.S., Smith A.M., Stecher T.P. Astron. J. 113, 1011 (1997) (From AAS CD-ROM series, Volume 8, 1997 Lee Brotzman [ADS] 09-Apr-1997) UIT: Ultraviolet Observations of the Small Magellanic Cloud

J/AJ/113/1026/  table1.dat table1.tex table2.dat table2.tex  (262Kb)

Gebhardt K., Pryor C., Williams T.B., Hesser J.E., Stetson P.B. Astron. J. 113, 1026 (1997) (From AAS CD-ROM series, Volume 8, 1997 Lee Brotzman [ADS] 09-Apr-1997) Fabry-Perot Observations of Globular Clusters. III. M15

J/AJ/113/1045/  table3  (14Kb)

Kozhurina-Platais V., Demarque P., Platais I., Orosz J.A., Barnes S. Astron. J. 113, 1045 (1997) The age of NGC 3680 and a test of convective overshoot.

J/AJ/113/1073/  table2 table2.tex table3 table3.tex  (548Kb)

Fry A.M., Carney B.W. Astron. J. 113, 1073 (1997) Chemical Abundances of Galactic Cepheid Variables That Calibrate the P-L Relation

J/AJ/113/1197/  doc.txt notes.dat table2.dat table3.dat  (144Kb)

Haynes M.P., Giovanelli R., Herter T., Vogt N.P., Freudling W., Maia M.A.G., Salzer J.J., Wegner G. Astron. J. 113, 1197 (1997) (From AAS CD-ROM series, Volume 8, 1997 Lee Brotzman [ADS] 02-May-1997) 21 cm H I Line Spectra of Galaxies in Nearby Clusters

J/AJ/113/1691/  table2.dat  (53Kb)

Hunter D.A., Vacca W.D., Massey P., Lynds R., O’Neil E.J. Astron. J. 113, 1691 (1997) (From AAS CD-ROM series, Volume 8, 1997 Lee Brotzman [ADS] 02-May-1997) Ultraviolet Photometry of Stars in the Compact Cluster R136 in the Large Magellanic Cloud

J/AJ/113/1711/  table8.dat  (53Kb)

Clement C.M., Shelton I. Astron. J. 113, 1711 (1997) (From AAS CD-ROM series, Volume 9, 1997 Lee Brotzman [ADS] 30-Oct-1997) The structure of the light curves of the RR Lyrae variables in the Oosterhoff type I cluster NGC 6171

J/AJ/113/1733/  table1.dat table3.dat  (209Kb)

Hillenbrand L.A. Astron. J. 113, 1733 (1997) On the stellar population and star-forming history of the Orion Nebula Cluster

J/AJ/113/1788/  table3.dat table3.tex  (75Kb)

Horner D.J., Lada E.A., Lada C.J. Astron. J. 113, 1788 (1997) (From AAS CD-ROM series, Volume 8, 1997 Lee Brotzman [ADS] 02-May-1997) A Near-Infrared Imaging Survey of NGC 2282

J/AJ/113/1939/  eps/* table2.dat table3.dat  (14.4 Mb)

Dickey J. M. Astron. J. 113, 1939 (1997) (From AAS CD-ROM series, Volume 9, 1997 Lee Brotzman [ADS] 29-Aug-1997) A VLA survey of the Hercules cluster. I. The H I data

J/AJ/113/2000/  table1 table1.tex  (29Kb)

Bischof O.B., Becker R.H. Astron. J. 113, 2000 (1997) Radio observations of 4079 quasars

J/AJ/113/2104/  table1 table2 table3  (12Kb)

Turner D.G., Mandushev G.I., Welch G.A. Astron. J. 113, 2104 (1997) Galactic clusters with associated Cepheid variables. V. The case of SU Cygni

J/AJ/113/2134 See IX/18/

J/AJ/114/122/  notes.dat table1.dat table1.tex table2.dat table2.tex  (160Kb)

Giovanelli R., Avera E., Karachentsev I.D. Astron. J. 114, 122 (1997) Spectroscopy of edge-on spirals.

J/AJ/114/189/  table3.dat  (84Kb)

Alcaino G., Liller W., Alvarado F., Kravtsov V., Ipatov A., Samus N., Smirnov O. Astron. J. 114, 189 (1997) The main sequence of the globular cluster M4=NGC 6121 from CCD NTT photometry.

J/AJ/114/198/  table1.dat table2.dat tables.tex  (151Kb)

Carpenter J.M., Meyer M.R., Dougados C., Strom S.E., Hillenbrand L.A. Astron. J. 114, 198 (1997) (From AAS CD-ROM series, Volume 9, 1997 Lee Brotzman [ADS] 26-Sep-1997) Properties of the Monoceros R2 Stellar Cluster

J/AJ/114/222/  table1 table2  (25Kb)

Weitenbeck A.J. Astron. J. 114, 222 (1997) Cross identifications in the region of cluster NGC 1502

J/AJ/114/326/  table1.dat table4.dat  (226Kb)

Alcock C., Allsman R.A., Alves D., Axelrod T.S., Becker A.C., Bennett D.P., Cook K.H., Freeman K.C., Griest K., Lacy C.H.S., Lehner M.J., Marshall S.L., Minniti D., Peterson B.A., Pratt M.R., Quinn P.J., Rodgers A.W., Stubbs C.W., Sutherland W., , Welch D.L. Astron. J. 114, 326 (1997) (From AAS CD-ROM series, Volume 9, 1997 Lee Brotzman [ADS] 05-Nov-1997) The MACHO Project LMC Variable Star Inventory. V. Classification and Orbits of 611 Eclipsing Binary Stars

J/AJ/114/341/  tables.dat tables.tex  (44Kb)

Gonzalez G., Lambert D.L. Astron. J. 114, 341 (1997) (From AAS CD-ROM series, Volume 9, 1997 Lee Brotzman [ADS] 26-Sep-1997) Abundance analyses of type II Cepheids in globular clusters

J/AJ/114/699/  table4.dat table6.dat  (77Kb)

Clampitt L., Burstein D. Astron. J. 114, 699 (1997) (From AAS CD-ROM series, Volume 9, 1997ASU] Oct-01-1997) Spectrophotometry of 237 stars in 7 open clusters

J/AJ/114/850/  table.dat  (75Kb)

Stone R.C. Astron. J. 114, 850 (1997) (From AAS CD-ROM series, Volume 9, 1997 Lee Brotzman [ADS] 27-Sep-1997) Systematic errors in the FK5 catalog as determined from CCD observations in the extragalactic reference frame

J/AJ/114/1087/  table1.dat table1.tex  (69Kb)

Mayor M., Meylan G., Udry S., Duquennoy A., Andersen J., Nordstrom B. Astron. J. 114, 1087 (1997) Radial velocities of southern stars obtained with the photoelectric scanner CORAVEL. VIII. Observations of 471 giant stars in omega Centauri.

J/AJ/114/1190/  table1.dat  (54Kb)

Garcia-Melendo E., Clement C.M.

Astron. J. 114, 1190 (1997) (From AAS CD-ROM series, Volume 9, 1997 Lee Brotzman [ADS] 31-Oct-1997) NSV 09295: A double-mode RR Lyrae variable

J/AJ/114/1458/  table2.dat table2.fit  (394Kb)

Mighell K.J. Astron. J. 114, 1458 (1997) (From AAS CD-ROM series, Volume 9, 1997 Lee Brotzman [ADS] 31-Oct-1997) WFPC2 observations of the Carina dwarf spheroidal galaxy

J/AJ/114/1471/  table3.dat table3.ps table3.tex  (495Kb)

Smith E.O., Michael R.R., Neill J.D. Astron. J. 114, 1471 (1997) (From AAS CD-ROM series, Volume 9, 1997 Lee Brotzman [ADS] 04-Nov-1997) Fornax globular cluster 3: new color-magnitude diagrams for cluster and surrounding field

J/AJ/114/1488/  table3.dat table4.dat table5.dat table6.dat tables.ps tables.tex  (4.5 Mb)

Holland S., Fahlman G.G., Richer H.B. Astron. J. 114, 1488 (1997) (From AAS CD-ROM series, Volume 9, 1997 Lee Brotzman [ADS] 04-Nov-1997) Deep HST V- and I-band observations of two globular clusters in the halo of M31

J/AJ/114/2269/  table1.dat table2.dat  (7Kb)

Kennefick J.D., Osmer P.S., Hall P.B., Green R.F. Astron. J. 114, 2269 (1997) A deep multicolor survey. III. Additional spectroscopy and implications for the number counts of faint quasars.

J/AJ/114/2353/  table3.dat  (21Kb)

Crampton D., Gussie G., Cowley A.P., Schmidtke P.C. Astron. J. 114, 2353 (1997) Probes for nearby galaxies.

J/AJ/114/2381/  tables.dat  (13Kb)

Miller B.W., Whitmore B.C., Schweizer F., Fall S.M. Astron. J. 114, 2381 (1997) The star cluster system of the merger remnant NGC 7252.

J/AJ/114/2402/  table3.dat table6.dat table7.dat tables.tex  (341Kb)

Courteau S. Astron. J. 114, 2402 (1997) Optical rotation curves and linewidths for Tully-Fisher applications.

J/AJ/114/2596/  table1.dat  (7Kb)

Blanco B.M., Blanco V.M. Astron. J. 114, 2596 (1997) RR Lyrae variables in a new galactic bulge window.

J/AJ/114/2644/  table1.dat  (67Kb)

Sung H., Bessell M.S., Lee S.-W. Astron. J., 114, 2644-2657 (1997) UBVRI H-alpha photometry of the young open cluster NGC 2264.

J/AJ/114/2822/  table.dat  (1.7 Mb)

Hutter D.J., Elias II N.M., Peterson E.R., Weaver W.B., Weaver G., Mozurkewich D., Vrba F.J., Simon R.S., Buscher D.F., Hummel C.A. Astron. J. 114, 2822 (1997) (From AAS CD-ROM series, Volume 9, 1997 Lee Brotzman [ADS] 04-Nov-1997) Seeing tests at four sites in support of the NPOI project

J/AJ/115/6/  refs.dat table1.dat table1.ps table1.tex table9.dat  (162Kb)

Barmby P., Huchra J.P. Astron. J. 115, 6 (1998) + Erratum vol. 116, p. 1508 Kinematics of the Hercules supercluster

J/AJ/115/130/  table3.dat table4.dat  (53Kb)

Sohn Y.J., Davidge T.J. Astron. J. 115, 130 (1998) VRI CCD photometry of supergiant stars in the barred galaxies NGC 925 and NGC 1637

J/AJ/115/168/  table1.dat  (44Kb)

Chiba M., Yoshii Y. Astron. J. 115, 168 (1998) Early evolution of the Galactic halo revealed from Hipparcos observations of metal-poor stars

J/AJ/115/193/  table1.dat table3.dat table4.dat  (60Kb)

Layden A.C. Astron. J. 115, 193 (1998)

RR Lyrae variables in the inner halo. I. Photometry.

J/AJ/115/296/  table1.dat table2.dat table3.dat  (142Kb)

Henden A.A., Stone R.C. Astron. J. 115, 296 (1998) New variables in the Sloan Digital Sky Survey calibration fields.

J/AJ/115/592/ table1.dat table2.dat table3.dat table4.dat table5.dat table6.dat table7.dat table8.dat (454Kb) Fischer P., Pryor C., Murray S., Mateo M., Richtler T. Astron. J. 115, 592 (1998) Mass segregation in young Large Magellanic Cloud clusters. I. NGC 2157.

J/AJ/115/635/  table2.dat table5.dat  (23Kb)

Metzger M.R., Caldwell J.A.R., Schechter P.L. Astron. J. 115, 635 (1998) The shape and scale of Galactic rotation from Cepheid kinematics

J/AJ/115/708/  non memb.dat table2.dat  (44Kb)

Drukier G. A., Slavin S. D., Cohn H. N., Lugger P. M., Berrington R. C., Murphy B. W., Seitzer P. O. Astron. J., 115, 708 (1998) The Global Kinematics of the Globular Clutser M15

J/AJ/115/734/  table2.dat table4.dat table5.dat  (104Kb)

Sung H., Bessell M.S., Lee S.-W. Astron. J. 115, 734 (1998) UBVRI and HαPhotometry of the young open cluster NGC 6231

J/AJ/115/1016/ table1.dat table2.dat table3.dat table4.dat table5.dat table6.dat table7.dat table8.dat (589Kb) Kaluzny J., Stanek K.Z., Krockenberger M., Sasselov D.D., Tonry J.L., Mateo M. Astron. J. 115, 1016 (1998) DIRECT distances to nearby galaxies using detached eclipsing binaries and Cepheids. I. Variables in the field M31B

J/AJ/115/1076/  table1.dat table5.dat  (42Kb)

Dawson D.W., Ianna P.A. Astron. J. 115, 1076 (1998) Collinder 110: an old open cluster in Monoceros.

J/AJ/115/1212 See I/275/

J/AJ/115/1253/  notes.dat table3.dat table4.dat table6.dat  (54Kb)

Perlman E.S., Padovani P., Giommi P., Sambruna R., Jones L.R., Tzioumis A., Reynolds J. Astron. J. 115, 1253 (1998) The Deep X-ray Radio Blazar Survey. I. Methods and first results.

J/AJ/115/1693 See VIII/65/

J/AJ/115/1801/  table2.dat  (8Kb)

Harris W.E., Harris G.L.H., McLaughlin D.E. Astron. J. 115, 1801 (1998) M87, globular clusters, and galactic winds: issues in giant galaxy formation.

J/AJ/115/1840/  field1.dat field2.dat field3.dat  (342Kb)

Hurley-Keller D., Mateo M., Nemec J. Astron. J. 115, 1840 (1998) The star formation history of the Carina dwarf galaxy.

J/AJ/115/1856/ table2.dat v1.dat v2.dat v3.dat v4.dat v5.dat v6.dat v7.dat v8.dat v9.dat v10.dat v11.dat v12.dat v13.dat v14.dat v15.dat v16.dat v17.dat v18.dat v19.dat v20.dat (42Kb) Mateo M., Hurley-Keller D., Nemec J. Astron. J. 115, 1856 (1998) Dwarf Cepheids in the Carina dwarf spheroidal galaxy.

J/AJ/115/1894/ table1.dat table2.dat table3.dat table4.dat table5.dat table6.dat table7.dat table8.dat (379Kb) Stanek K.Z., Kaluzny J., Krockenberger M., Sasselov D.D., Tonry J.L., Mateo M. Astron. J. 115, 1894 (1998) DIRECT distances to nearby galaxies using detached eclipsing binaries and Cepheids. II. Variables in the field M31A.

J/AJ/115/1989/  table3.dat  (102Kb)

Tajitsu A., Tamura S. Astron. J., 115, 1989-2008 (1998) A new distance indicator to galactic planetary nebulae based upon IRAS fluxes.

J/AJ/115/2397/ note1.dat table1.dat table2.dat table3.dat table4.dat table5.dat table6.dat table8.dat table10.dat table11.dat table12.dat table14.dat table15.dat table17.dat table18.dat (148Kb) Eggen O.J. Astron. J. 115, 2397 (1998) Kinematics and metallicity of stars in the solar region.

J/AJ/115/2587/  table1.dat table2.dat  (85Kb)

Thompson R.J., Shelton R.G., Arning C.A. Astron. J. 115, 2587 (1998) Initial results of a comprehensive ultrasoft survey of the Einstein IPC database: source list and confirmation of the selection procedure.

J/AJ/116/1/  refs.dat table2.dat table3.dat  (400Kb)

Da Costa L.N., Willmer C.N.A., Pellegrini P.S., Chaves O.L., Rite C., Maia M.A.G., Geller M.J., Latham D.W., Kurtz M.J., Huchra J.P., Ramella M., Fairall A.P., Smith C., Lipari S. Astron. J. 116, 1 (1998) The Southern Sky Redshift Survey.

J/AJ/116/102/  table1.dat  (11Kb)

Tyson J.A., Fischer P., Guhathakurta P., McIlroy P., Wenk R., Huchra J., Macri L., Neuschaefer L., Sarajedini V., Glazebrook K., Ratnatunga K., Griffiths R. Astron. J. 116, 102 (1998) Deep optical imaging of the bright Seyfert galaxy NGC 5548: a long, very low surface brightness tail.

J/AJ/116/146/  table2.dat  (2.0 Mb)

Lynds R., Tolstoy E., O’Neil E.J.Jr, Hunter D.A. Astron. J. 116, 146 (1998) Star formation in and evolution of the blue compact dwarf galaxy UGC 6456 determined from Hubble space telescope images

J/AJ/116/180/  field.dat table2.dat  (3.1 Mb)

Parker J.W., Hill J.K., Cornett R.H., Hollis J., Zamkoff E., Bohlin R.C., O’Connell R.W., Neff S.G., Roberts M.S., Smith A.M., Stecher T.P. Astron. J. 116, 180 (1998) Ultraviolet Imaging Telescope observations of the Magellanic clouds

J/AJ/116/469/  table1.dat table3.dat table4.dat  (44Kb)

Metzger M.R., Schechter P.L. Astron. J. 116, 469 (1998) A search for distant galactic Cepheids toward l = 60^{∘}.

J/AJ/116/516 See I/251/

J/AJ/116/560/  table2.dat table3.dat  (30Kb)

Postman M., Lubin L.M., Oke J.B. Astron. J. 116, 560 (1998) A study of nine high-redshift clusters of galaxies. II. Photometry, spectra, and ages of clusters 0023+0423 and 1604+4304.

J/AJ/116/584/  table1a.dat table1b.dat table2a.dat table2b.dat table4a.dat table4b.dat  (246Kb)

Lubin L.M., Postman M., Oke J.B., Ratnatunga K.U., Gunn J.E., Hoessel J.G., Schneider D.P. Astron. J. 116, 584 (1998) A study of nine high-redshift clusters of galaxies. III. Hubble space telescope morphology of clusters 0023+0423 and 1604+4304.

J/AJ/116/723/  tab134.dat table2.dat  (144Kb)

Bica E., Geisler D., Dottori H., Claria J.J., Piatti A.E., Santos J.F.C.Jr Astron. J. 116, 723 (1998) Ages and metallicities of star clusters and surrounding fields in the outer disk of the Large Magellanic Cloud.

J/AJ/116/917/  table1.dat  (34Kb)

Worley C.E., Mason B.D. Astron. J. 116, 917 (1998) Micrometer measures of double stars

J/AJ/116/941/  table3.dat table4.dat  (27Kb)

Brandt J.C., Heap S.R., Beaver E.A., Boggess A., Carpenter K.G., Ebbets D.C., Hutchings J.B., Jura M., Leckrone D.S., Linsky J.L., Maran S.P., Savage B.D., Smith A.M., Trafton L.M., Walter F.M., Weymann R.J., Snow M., Ake T.B., Hogen R.H. Astron. J. 116, 941 (1998) An ultraviolet spectral atlas of 10 Lacertae obtained with the Goddard high resolution spectrograph on the Hubble Space Telescope.

J/AJ/116/1094/  table2.dat table3.dat  (246Kb)

Cabanela J.E., Aldering G. Astron. J. 116, 1094 (1998) Galaxy alignments in the Pisces-Perseus supercluster revisited.

J/AJ/116/1275/ table2a.dat table2b.dat table2c.dat table2d.dat table2e.dat table2f.dat table2g.dat (8.3 Mb) Dolphin A.E., Hunter D.A. Astron. J. 116, 1275 (1998) Recent star formation in Shapley constellation III in the Large Magellanic Cloud

J/AJ/116/1447/  tablea1.dat tablea2.dat  (54Kb)

Varricatt W.P., Ashok N.M., Chandrasekhar T. Astron. J. 116, 1447 (1998) Near-infrared photometric studies of RZ Cassiopeiae.

J/AJ/116/1529/ a779.dat a1691.dat a1749.dat a1767.dat a1837.dat a1927.dat a2067.dat a2079.dat a2089.dat a2199.dat a2666.dat (56Kb) Hill J.M., Oegerle W.R. Astron. J. 116, 1529 (1998) Dynamics of cD clusters of galaxies. III. Redshift data for 11 Abell clusters.

J/AJ/116/1573/  table3.dat table3.ps table3.tex  (159Kb)

Barton E.J., De Carvalho R.R., Geller M.J. Astron. J. 116, 1573 (1998) Environments of Redshift Survey Compact Groups of galaxies.

J/AJ/116/1626/  table1.dat table3.dat  (114Kb)

Baggett W.E., Baggett S.M., Anderson K.S.J. Astron. J. 116, 1626 (1998) Bulge-disk decomposition of 659 spiral and lenticular galaxy brightness profiles

J/AJ/116/1757/  table1.dat  (434Kb)

Guhathakurta P., Webster Z.T., Yanny B., Schneider D.P., Bahcall J.N. Astron. J. 116, 1757 (1998) Globular cluster photometry with the Hubble Space Telescope. VII. Color gradients and blue stragglers in the central region of M30 from Wide Field Planetary Camera 2 observations.

J/AJ/116/1775/  table2.dat table3.dat table4.dat table7.dat  (846Kb)

Rey S.-C., Lee Y.-W., Byun Y.-I., Chun M.-S. Astron. J. 116, 1775 (1998) CCD photometry of the globular cluster M 53. I. Color-magnitude data and blue straggler stars.

J/AJ/116/1801/  table2.dat  (31Kb)

Delgado A.J., Alfaro E.J., Moitinho A., Franco J. Astron. J. 116, 1801 (1998) Pre-main-sequence stars in the young galactic cluster IC 4996: a CCD photometric study.

J/AJ/116/1816/  table1.dat  (138Kb)

Hillenbrand L.A., Strom S.E., Calvet N., Merrill K.M., Gatley I., Makidon R.B., Meyer M.R., Skrutskie M.F. Astron. J. 116, 1816 (1998) Circumstellar disks in the Orion nebula cluster.

J/AJ/116/1922/  table1.dat  (29Kb)

Anthony-Twarog B.J., Twarog B.A. Astron. J. 116, 1922 (1998) Ca II H and K photometry on the uvby system. III. The metallicity calibration for the red giants.

J/AJ/116/1998/  orbits.dat  (11Kb)

Levison H.F., Lissauer J.J., Duncan M.J. Astron. J. 116, 1998 (1998) Modeling the diversity of outer planetary systems.

J/AJ/116/2081/  table2.dat  (58Kb)

Wang Y., Bahcall N., Turner E.L.

Astron. J. 116, 2081 (1998) A catalog of color-based redshift estimates for z 4 galaxies in the Hubble Deep Field.

J/AJ/116/2108/  table1.dat  (15Kb)

Koranyi D.M., Geller M.J., Mohr J.J., Wegner G. Astron. J. 116, 2108 (1998) A photometric and kinematic study of AWM 7.

J/AJ/116/2246/  table3.dat table4.dat table5.dat table6.dat  (20Kb)

Kingsburgh R.L., McCall M.L. Astron. J. 116, 2246 (1998) Automatic determination of unbiased luminosity functions for H II regions. II. Four nearby dwarf galaxies.

J/AJ/116/2341/  sextansa.dat  (293Kb)

Van Dyk S.D., Puche D., Wong T. Astron. J. 116, 2341 (1998) Recent star formation in Sextans A.

J/AJ/116/2395/  tables.dat  (1.6 Mb)

Mighell K.J., Sarajedini A., French R.S. Astron. J. 116, 2395 (1998) WFPC2 observations of star clusters in the Magellanic clouds. II. The oldest star clusters in the Small Magellanic Cloud.

J/AJ/116/2415/  photom.dat  (248Kb)

Alcaino G., Liller W., Alvarado F., Kravtsov V., Ipatov A., Samus N., Smirnov O. Astron. J. 116, 2415 (1998) Multicolor CCD photometry of the poorly studied globular cluster M80.

J/AJ/116/2549/  table1.dat  (22Kb)

Samec R.G., Carrigan B.J., McDermith R.J. Astron. J. 116, 2549 (1998) HW Persei: an eclipsing binary at critical contact?

J/AJ/116/2556 See I/277/

J/AJ/116/2728/  table3.dat  (89Kb)

Scodeggio M., Giovanelli R., Haynes M.P. Astron. J. 116, 2728 (1998) The universality of the fundamental plane of E and S0 galaxies: sample definition and I-band photometric data.

J/AJ/116/2738/  table2.dat table3.dat  (42Kb)

Scodeggio M., Giovanelli R., Haynes M.P. Astron. J. 116, 2738 (1998) The universality of the fundamental plane of E and S0 galaxies: spectroscopic data.

J/AJ/116/2805/  table1.dat table2.dat table3.dat table4.dat table5.dat table6.dat  (89Kb)

Van Zee L., Salzer J.J., Haynes M.P., O’Donoghue A.A., Balonek T.J. Astron. J. 116, 2805 (1998) Spectroscopy of outlying H II regions in spiral galaxies: abundances and radial gradients.

J/AJ/117/17/  table4.dat table6.dat  (77Kb)

Thompson R.I., Storrie-Lombardi L.J., Weymann R.J., Rieke M.J., Schneider G., Stobie E., Lytle D. Astron. J. 117, 17 (1999) Near-infrared camera and multi-object spectrometer observations of the Hubble Deep Field: observations, data reduction, and galaxy photometry.

J/AJ/117/75/  table2.dat table3.dat table4.dat table5.dat table6.dat table10.dat  (28Kb)

Conselice C.J., Gallagher J.S.III Astron. J. 117, 75 (1999) Tides, interactions, and fine-scale substructures in galaxy clusters.

J/AJ/117/206/  table1.dat table2.dat table3.dat table4.dat table5.dat  (47Kb)

Brosch N., Shara M., Mackenty J., Zurek D., McLean B. Astron. J. 117, 206 (1999) Far-ultraviolet imaging of the field star population in the Large Magellanic Cloud with the Hubble Space Telescope.

J/AJ/117/238/  table1.dat table2.dat  (569Kb)

Bica E.L.D., Schmitt H.R., Dutra C.M., Oliveira H.L. Astron. J. 117, 238 (1999)

A revised and extended catalog of Magellanic system clusters, associations, and emission nebulae. II. The Large Magellanic Cloud.

J/AJ/117/247/  table2.dat  (19Kb)

Stetson P.B., Bolte M., Harris W.E., Hesser J.E., Van den Bergh S., Vandenberg D.A., Bell R.A., Johnson J.A., Bond H.E., Fullton L.K., Fahlamn G.G., Richer H.B. Astron. J. 117, 247 (1999) Ages for globular clusters in the outer galactic halo: the second-parameter clusters Palomar 3, Palomar 4, and Eridanus

J/AJ/117/286/  table2.dat table3.dat  (27Kb)

Wehlau A., Slawson R. W., Nemec J. M. Astron. J. 117, 286 (1999) Two-color CCD photometry of variable stars in NGC 7006

J/AJ/117/354/  tablec1.dat  (23Kb)

De Zeeuw P.T., Hoogerwerf R., De Bruijne J.H.J., Brown A.G.A., Blaauw A. Astron. J. 117, 354 (1999) A Hipparcos census of the nearby OB associations.

J/AJ/117/446/  table1.dat table2.dat  (13Kb)

Chen Y., Yao Y., Yang J., Hirao T., Ishii M., Nagata T., Sato S. Astron. J. 117, 446 (1999) Near-infrared imaging of the star-forming region AFGL 5157.

J/AJ/117/548/  notes.dat table2.dat table3.dat table4.dat table6.dat  (37Kb)

Horch E., Ninkov Z., Van Altena W.F., Meyer R.D., Girard T.M., Timothy J.G. Astron. J. 117, 548 (1999) Speckle observations of binary stars with the WIYN telescope. I. Measures during 1997.

J/AJ/117/587/  refs.dat table1.dat  (80Kb)

Petrova A.V., Orlov V.V. Astron. J. 117, 587 (1999) Apsidal motion in double stars. I. Catalog.

J/AJ/117/981/  refs.dat table4.dat table6.dat  (154Kb)

Beers T.C., Rossi S., Norris J.E., Ryan S.G., Shefler T. Astron. J. 117, 981 (1999) Estimation of stellar metal abundance. II. A recalibration of the Ca II K technique, and the autocorrelation function method.

J/AJ/117/1313/  notes.dat table1.dat table6.dat var.dat  (814Kb)

Layden A.C., Ritter L.A., Welch D.L., Webb T.M.A. Astron. J. 117, 1313 (1999) The variable stars and blue horizontal branch of the metal-rich globular cluster NGC 6441.

J/AJ/117/1375/  table1.dat  (24Kb)

Simon M., Close L.M., Beck T.L. Astron. J. 117, 1375 (1999) Adaptive optics imaging of the Orion Trapezium cluster.

J/AJ/117/1700/  table1.dat  (36Kb)

Carlson M.N., Holtzman J.A., Grillmair C.J., Mould J.R., Griffiths R.E., Ballester G.E., Burrows C.J., Clarke J.T., Crisp D., Evans R.W., Gallagher III J.S., Hester J.J., Hoessel J.G., Scowen P.A., Stapelfeldt K.R., Trauger J.T., Watson A.M., Westphal J.A. Astron. J. 117, 1700 (1999) Deep Hubble Space Telescope observations of blue star clusters in NGC 3597

J/AJ/117/1708/  table2.dat  (45Kb)

Johnson K.E., Vacca W.D., Leitherer C., Conti P.S., Lipscy S.J. Astron. J. 117, 1708 (1999) The very young starburst merger system NGC 1741.

J/AJ/117/1890/  notes.dat table2.dat table5.dat table6.dat table7.dat  (40Kb)

Mason B.D., Martin C., Hartkopf W.I., Barry D.J., Germain M.E., Douglass G.G., Worley C.E., Wycoff G.L., Ten Brummelaar T., Franz O.G. Astron. J. 117, 1890 (1999) Speckle interferometry of new and problem Hipparcos binaries.

J/AJ/117/1905/  table2.dat  (224Kb)

Germain M.E., Douglass G.G., Worley C.E.

Astron. J. 117, 1905 (1999) Speckle interferometry at the US naval observatory. II

J/AJ/117/2039/  notes.dat table1.dat table3.dat  (151Kb)

Haynes M.P., Giovanelli R., Chamaraux P., Da Costa L.N., Freudling W., Salzer J.J., Wegner G. Astron. J. 117, 2039 (1999) The I-band Tully-Fisher relation for Sc galaxies: 21 centimeter H I line data.

J/AJ/117/2244/  table1a.dat table1b.dat table1c.dat table1d.dat table1e.dat table1f.dat  (11.3 Mb)

Olsen K.A.G. Astron. J. 117, 2244 (1999) Star formation histories from Hubble space telescope color-magnitude diagrams of six fields of the Large Magellanic Cloud

J/AJ/117/2308/  table1.dat  (67Kb)

Wilhelm R., Beers T.C., Gray R.O. Astron. J. 117, 2308 (1999) Spectroscopy of hot stars in the galactic halo. II. The identification and classification of horizontal-branch and other A-type stars

J/AJ/117/2329/ table1.dat table2a.dat table2b.dat table2c.dat table2d.dat table3.dat tablea1.dat tableb1.dat (223Kb) Wilhelm R., Beers T.C., Sommer-Larsen J., Pier J.R., Layden A.C., Flynn C., Rossi S., Christensen P.R. Astron. J. 117, 2329 (1999) Spectroscopy of hot stars in the galactic halo. III. Analysis of a large sample of field horizontal-branch and other A-type stars

J/AJ/117/2466 See IX/35/

J/AJ/117/2511/  table4.dat  (165Kb)

Germain M.E., Douglass G.G., Worley C.E. Astron. J. 117, 2511 (1999) Speckle interferometry at the US Naval Observatory. III.

J/AJ/117/2757/  table5.dat  (30Kb)

Zheng Z., Shang Z., Su H., Burstein D., Chen J., Deng Z., Byun Y.-I., Chen R., Chen W.-P., Deng L., Fan X., Fang Li-Z., Hester J.J., Jiang Z., Li Y., Lin W., Sun W.-H., Tsay W.-S., Windhorst R.A., Wu H., Xia X., Xu W., Xue S., Yan H., Zheng Z., Zhou X., Zhu J., Zou Z., Lu P. Astron. J. 117, 2757 (1999) Deep intermediate-band surface photometry of NGC 5907

J/AJ/117/2810/ table1.dat table2.dat table3.dat table4.dat table5.dat table6.dat table7.dat table8.dat (1.0 Mb) Stanek K.Z., Kaluzny J., Krockenberger M., Sasselov D.D., Tonry J.L., Mateo M. Astron. J. 117, 2810 (1999) DIRECT distances to nearby galaxies using detached eclipsing binaries and Cepheids. III. Variables in the field M31C.

J/AJ/117/2868/  table2.dat table3.dat table4.dat  (950Kb)

Lee J.-W., Carney B.W. Astron. J. 117, 2868 (1999) BV photometry of RR Lyrae variables in the globular cluster M2 (NGC 7089).

J/AJ/117/2895 See I/265/

J/AJ/117/2941/  table1.dat table2.dat  (33Kb)

Stassun K.G., Mathieu R.D., Mazeh T., Vrba F.J. Astron. J. 117, 2941 (1999) The rotation period distribution of pre-main-sequence stars in and around the Orion nebula.

J/AJ/117/2980/  table12.dat table13.dat  (62Kb)

Varricatt W.P., Ashok N.M. Astron. J. 117, 2980 (1999) Near-infrared photometric studies of R Canis Majoris

J/AJ/118/346/ table1.dat table2.dat table3.dat table4.dat table5.dat table6.dat table7.dat table8.dat (27Kb) Kaluzny J., Mochejska B.J., Stanek K.Z., Krockenberger M., Sasselov D.D., Tonry J.L., Mateo M. Astron. J. 118, 346 (1999) DIRECT distances to nearby galaxies using detached eclipsing binaries and Cepheids. IV. Variables in the field M31D.

J/AJ/118/625/  table2.dat table3.dat  (12Kb)

Dwarakanath K.S., Owen F.N. Astron. J. 118, 625 (1999) On the different radio source populations in the Butcher-Oemler clusters Abell 2125 and 2645.

J/AJ/118/895/  table1.dat table3.dat  (71Kb)

Tiede G.P., Terndrup D.M. Astron. J. 118, 895 (1999) Kinematics, metallicities, and stellar distributions in the inner disk and bulge of the Milky way.

J/AJ/118/1131/  table4.dat table6.dat  (32Kb)

Willmer C.N.A., Maia M.A.G., Mendes S.O., Alonso M.V., Rios L.A., Chaves O.L., De Mello D.F. Astron. J. 118, 1131 (1999) Two galaxy clusters: A3565 and A3560

J/AJ/118/1245/  table2.dat  (17Kb)

Brown J.A., Wallerstein G., Gonzalez G. Astron. J. 118, 1245 (1999) Elemental abundances in five stars in M54, a globular cluster associated with the Sagittarius galaxy

J/AJ/118/1395/  table2.dat  (86Kb)

Douglass G.G., Mason B.D., Germain M.E., Worley C.E. Astron. J. 118, 1395 (1999) Speckle interferometry at the US naval observatory. IV.

J/AJ/118/1435/  table2.dat  (26Kb)

Wadadekar Y., Kembhavi A. Astron. J. 118, 1435 (1999) A study of quasar radio emission from the VLA FIRST survey

J/AJ/118/1468/  notes.dat table3.dat table4.dat table5.dat  (80Kb)

Dale D.A., Giovanelli R., Haynes M.P., Hardy E., Campusano L.E. Astron. J. 118, 1468 (1999) Seeking the local convergence depth. IV. Tully-Fisher observations of 35 Abell clusters.

J/AJ/118/1671/  for4.dat  (891Kb)

Buonanno R., Corsi C.E., Castellani M., Marconi G., Fusi Pecci F., Zinn R. Astron. J. 118, 1671 (1999) Hubble space telescope photometry of the Fornax dwarf spheroidal galaxy: cluster 4 and its field.

J/AJ/118/2014/  table2.dat table3.dat table4.dat  (277Kb)

White R.A., Bliton M., Bhavsar S.P., Bornmann P., Burns J.O., Ledlow M.J., Loken C. Astron. J. 118, 2014 (1999) A catalog of nearby poor clusters of galaxies

J/AJ/118/2071/  table1.dat table2.dat  (54Kb)

Buta R., Crocker D.A., Byrd G.G. Astron. J. 118, 2071 (1999) A Hubble Space Telescope optical and ground-based near-infrared study of the giant nuclear ring in ESO 565-11.

J/AJ/118/2211/ table1.dat table2.dat table3.dat table4.dat table5.dat table6.dat table7.dat table8.dat table9.dat (443Kb) Mochejska B.J., Kaluzny J., Stanek K.Z., Krockenberger M., Sasselov D.D. Astron. J. 118, 2211 (1999) DIRECT distances to nearby galaxies using detached eclipsing binaries and Cepheids. V. Variables in the field M31F.

J/AJ/118/2409/  table1.dat  (11Kb)

Dolan C.J., Mathieu R.D. Astron. J. 118, 2409 (1999) A WIYN lithium survey for young stars in the λOrionis star-forming region.

J/AJ/118/2488 See I/263/

J/AJ/118/2561/  table2.dat  (26Kb)

Grogin N.A, Geller M.J. Astron. J. 118, 2561 (1999) An imaging and spectroscopic survey of galaxies within prominent nearby voids. I. The sample and luminosity distribution

J/AJ/118/2839/  table2.dat table3.dat  (804Kb)

Testa V., Ferraro F.R., Chieffi A., Straniero O., Limongi M., Fusi Pecci F. Astron. J. 118, 2839 (1999) The Large Magellanic Cloud globular cluster NGC 1866: new data, new models, new analysis.

J/AJ/119/32/  table1.dat table2.dat  (50Kb)

Grogin N.A., Geller M.J. Astron. J. 119, 32 (2000) An imaging and spectroscopic survey of galaxies within prominent nearby voids. II. Morphologies, star formation, and faint companions.

J/AJ/119/44/  table1.dat  (21Kb)

Koranyi D.M., Geller M.J. Astron. J. 119, 44 (2000) Kinematics and mass profile of AWM 7

J/AJ/119/486/  table2.dat  (69Kb)

Gardner J.P., Baum S.A., Brown T.M., Carollo C.M., Christensen J., Dashevsky I., Dickinson M.E., Espey B.R., Ferguson H.C., Fruchter A.S., Gonnella A.M., Gonzalez-Lopezlira R.A., Hook R.N., Kaiser M.E., Martin C.L., Sahu K.C., Savaglio S., Smith T.E., Teplitz H.I., Williams R.E., Wilson J. Astron. J. 119, 486 (2000) The Hubble Deep Field South: STIS imaging.

J/AJ/119/727/  refs.dat table1.dat table2.dat table3.dat table5.dat  (81Kb)

Barmby P., Huchra J.P., Brodie J.P., Forbes D.A., Schroder L.L., Grillmair C.J. Astron. J. 119, 727 (2000) M31 globular clusters: colors and metallicities

J/AJ/119/777/  table2.dat  (8Kb)

Lee M.G., Kim S.C. Astron. J. 119, 777 (2000) Stellar populations of the Sagittarius dwarf irregular galaxy

J/AJ/119/923 See II/226/

J/AJ/119/936/  table3.dat  (18Kb)

McDonald S.W., Elliot J.L. Astron. J. 119, 936 (2000) Triton stellar occultation candidates: 2000-2009

J/AJ/119/1214/  table1.dat table3.dat  (117Kb)

Bica E., Dutra C.M. Astron. J. 119, 1214 (2000) Updating the census of star clusters in the Small Magellanic Cloud

J/AJ/119/1225/  table2.dat table3.dat table4.dat  (280Kb)

Hughes J. Wallerstein G. Astron. J. 119, 1225 (2000) Age and metallicity effects in Omega Centauri: Stroemgren photometry at the main-sequence turnoff

J/AJ/119/1268/  table1.dat table2.dat  (130Kb)

Gebhardt K., Pryor C., O’Connell R.D., Williams T.B., Hesser J.E. Astron. J. 119, 1268 (2000) Canada-France-Hawaii telescope adaptive optics observations of the central kinematics in M15

J/AJ/119/1389/  table6.dat table7.dat  (18Kb)

Lacy C.H.S., Torres G., Claret A., Stefanik R.P., Latham D.W., Sabby J.A. Astron. J. 119, 1389 (2000) Absolute properties of the eclipsing binary star FS Monocerotis

J/AJ/119/1424/  table5.dat  (33Kb)

Houdashelt M.L., Bell R.A., Sweigart A.V., Wing R.F. Astron. J. 119, 1424 (2000) Synthetic spectra and color-temperature relations of M giants

J/AJ/119/1448/  table2.dat table5.dat  (96Kb)

Houdashelt M.L., Bell R.A., Sweigart A.V. Astron. J. 119, 1448 (2000) Improved color-temperature relations and bolometric corrections for cool stars

J/AJ/119/1562/  table3.dat  (1.4 Mb)

Margoniner V. E., de Carvalho R. R. Astron. J. 119, 1562 (2000)

Photometric properties of 48 clusters of galaxies. I. The Butcher-Oemler effect

J/AJ/119/1748/ table2.dat table3.dat table4.dat table5.dat table7a.dat table7b.dat table7c.dat (540Kb) Keller S.C., Bessell M.S., Da Costa G.S. Astron. J. 119, 1748 (2000) Wide field planetary camera 2 imaging of young clusters in the Magellanic clouds.

J/AJ/119/1760/ notes.dat table2.dat table3.dat table4.dat table5.dat table6.dat table8.dat table9.dat table10.dat table11.dat (3.7 Mb) Layden A.C., Sarajedini A. Astron. J. 119, 1760 (2000) Photometry of the globular cluster M54 and the Sagittarius Dwarf Galaxy: the age-metallicity relation.

J/AJ/119/1824/  table4.dat table5.dat  (35Kb)

Rey S.-C., Lee Y.-W., Joo J.-M., Walker A., Baird S. Astron. J. 119, 1824 (2000) CCD photometry of the globular cluster ωCentauri. I. Metallicity of RR Lyrae stars from caby photometry

J/AJ/119/1901/  table1.dat  (178Kb)

Akerlof C., Amrose S., Balsano R., Bloch J., Casperson D., Fletcher S., Gisle R G., Hills J., Kehoe R., Lee B., Marshall S., McKay T., Pawl A., Schaefer J., Szymanski J.A., Wren J. Astron. J. 119, 1901 (2000) ROTSE all-sky surveys for variable stars. I. Test fields

J/AJ/119/2008/  table4.dat  (16Kb)

Stone R.C., McDonald S.W., Elliot J.L., Bowell E. Astron. J. 119, 2008 (2000) 10199 Chariklo stellar occultation candidates: 1999-2005.

J/AJ/119/2214/  table2.dat table5.dat  (90Kb)

Massey P., Waterhouse E., DeGioia-Eastwood K. Astron. J. 119, 2214 (2000) The progenitor masses of Wolf-Rayet stars and luminous blue variables determined from cluster turnoffs. I. Results from 19 OB Associations in the Magellanic Clouds

J/AJ/119/2274/  table2.dat  (90Kb)

Ivanov V.D., Borissova J., Alonso-Herrero A., Russeva T. Astron. J. 119, 2274 (2000) Extending the red giant branch versus metallicity calibration toward metal-poor systems: near-infrared photometry of the galactic globular clusters M56 and M15

J/AJ/119/2282/  table3.dat  (51Kb)

Anthony-Twarog B.J., Twarog B. A. Astron. J. 119, 2282 (2000) CCD uvbyHβphotometry in clusters. I. The open cluster standard, IC 4651

J/AJ/119/2540/  table2.dat table3.dat table4.dat  (39Kb)

Grazian A., Cristiani S., D’Odorico V., Omizzolo A., Pizzella A. Astron. J. 119, 2540 (2000) The Asiago-ESO/RASS QSO survey. I. The catalog and the local QSO luminosity function.

J/AJ/119/2571/  phothdf.dat table2.dat table4.dat table5.dat table6.dat  (1.1 Mb)

Vanden Berk D.E., Stoughton C., Crotts A.P.S., Tytler D., Kirkman D. Astron. J. 119, 2571 (2000) QSOs and absorption-line systems surrounding the Hubble Deep Field.

J/AJ/119/2598/  table3.dat  (257Kb)

Sowards-Emmerd D., Smith J.A., McKay T.A., Sheldon E., Tucker D.L., Castander F.J. Astron. J. 119, 2598 (2000) A catalog of photometry for Las Campanas Redshift Survey galaxies on the Sloan Digital Sky Survey system.

J/AJ/119/2780/  table2.dat  (13Kb)

Albert L., Demers S., Kunkel W.E. Astron. J. 119, 2780 (2000) A carbon star survey of the Local Group dwarf galaxies. I. IC 1613.

J/AJ/119/2801/  s8.dat s14.dat  (64Kb)

Langston G., Minter A., D’Addario L., Eberhardt K., Koski K., Zuber J. Astron. J. 119, 2801 (2000) The first galactic plane survey at 8.35 and 14.35 GHz.

J/AJ/119/2866/  refs.dat table1.dat table3.dat table4.dat  (703Kb)

Beers T.C., Chiba M., Yoshii Y., Platais I., Hanson R.B., Fuchs B., Rossi S. Astron. J. 119, 2866 (2000) Kinematics of metal-poor stars in the Galaxy. II. Proper motions for a large non-kinematically selected sample.

J/AJ/119/3026/  table1.dat table2.dat table3.dat  (154Kb)

Rebull L.M., Hillenbrand L.A., Strom S.E., Duncan D.K., Patten B.M., Pavlovsky C.M., Makidon R., Adams M.T. Astron. J. 119, 3026 (2000) Circumstellar disk candidates identified from ultraviolet excesses in the Orion nebula cluster flanking fields.

J/AJ/119/3071/  table3.dat  (102Kb)

Douglass G.G., Mason B.D., Rafferty T.J., Holdenried E.R., Germain M.E. Astron. J. 119, 3071 (2000) Speckle interferometry at the US naval observatory. V.

J/AJ/119/3084/  notes.dat position.dat table2.dat table4.dat table5.dat  (256Kb)

Hartkopf W.I., Mason B.D., McAlister H.A., Roberts L.C.Jr, Turner N.H., Ten Brummelaar T.A., Prieto C.M., Ling J.F., Franz O.G. Astron. J. 119, 3084 (2000) ICCD speckle observations of binary stars. XXIII. Measurements during 1982-1997 from six telescopes, with 14 new orbits.

J/AJ/120/1/  tables.dat  (47Kb)

Adami C., Holden B.P., Castander F.J., Nichol R.C., Mazure A., Ulmer M.P., Postman M., Lubin L.M. Astron. J. 120, 1 (2000) The Canada-France-Hawaii telescope optical PDCS survey (COP). I. The data.

J/AJ/120/284/  table1.dat table2.dat table4.dat table5.dat  (352Kb)

Siegel M.H., Majewski S.R. Astron. J. 120, 284 (2000) Exploring the Leo II dwarf spheroidal galaxy. I. The variable star content.

J/AJ/120/333/  table1.dat table3.dat table4.dat table5.dat  (116Kb)

Sung H., Chun M.-Y., Bessell M.S. Astron. J. 120, 333 (2000) UBVRI and Hαphotometry of the young open cluster NGC 6530.

J/AJ/120/1120/  notes.dat refs.dat table1.dat table2.dat  (64Kb)

Mason B.D., Hartkopf W.I., Holdenried E.R., Rafferty T.J., Wycoff G.L., Hennessy G.S., Hall D.M., Urban S.E., Corbin T.E. Astron. J. 120, 1120 (2000) Speckle interferometry at the US Naval Observatory. VI.

J/AJ/120/1306/  table2.dat  (13Kb)

Lelievre M., Roy J.R. Astron. J. 120, 1306 (2000) The H II regions of the extreme outer disk of NGC 628

J/AJ/120/1364/  stars.dat table2.dat table9.dat  (20Kb)

Cavallo R.M., Nagar N.M. Astron. J. 120, 1364 (2000) Aluminum abundances, deep mixing, and the blue-tail second-parameter effect in the globular clusters M3 and M13

J/AJ/120/1426/  notes.dat table1.dat table2.dat  (37Kb)

Garmire G., Feigelson E.D., Broos P., Hillenbrand L.A., Pravdo S.H., Townsley L., Tsuboi Y. Astron. J. 120, 1426 (2000) Chandra X-ray observatory study of the Orion nebula cluster and BN/KL region.

J/AJ/120/1548/  table2.dat table3.dat  (22Kb)

Elias II N.M., Koch R.H. Astron. J. 120, 1548 (2000) Photospheric spots and a chromospheric plage on V523 Cassiopeiae

J/AJ/120/1801/  table4.dat table5.dat  (9Kb)

Battinelli P., Demers S. Astron. J. 120, 1801 (2000) A carbon star survey of the Local Group dwarf galaxies. II. Pegasus, DDO 210, and Tucana.

J/AJ/120/1808/  table1.dat table6.dat  (1.2 Mb)

Cole A.A., Smecker-Hane T.A., Gallagher III J.S.

Astron. J. 120, 1808 (2000) The metallicity distribution function of red giants in the Large Magellanic Cloud.

J/AJ/120/1841/  ew.dat stars.dat  (989Kb)

Fulbright J.P. Astron. J., 120, 1841 (2000) Abundances and Kinematics of Field Halo and Disk Stars. I. Observational Data and Abundance Analysis

J/AJ/120/2101/  table1.dat table2.dat table3.dat  (22Kb)

Marchenko S.V., Arias J., Barba R., Balona L., Moffat A.F.J., Niemela V.S., Shara M.M., Sterken C. Astron. J., 120, 2101-2113 (2000) The puzzle of HD 104994 (WR 46).

J/AJ/120/2131 See I/268/

J/AJ/120/2190/  table1.dat table2.dat table3.dat  (22Kb)

van den Bergh S., Cohen J.G., Hogg D.W., Blandford R. Astron. J. 120, 2190 (2000) Caltech Faint Galaxy Redshift Survey. XIV. Galaxy morphology in the Hubble Deep Field (North) and its flanking fields to z=1.2.

J/AJ/120/2206/  table2.dat table2.tex  (81Kb)

Fontana A., D’Odorico S., Poli F., Giallongo E., Arnouts S., Cristiani S., Moorwood A., Saracco P. Astron. J. 120, 2206 (2000) Photometric redshifts and selection of high-redshift galaxies in the NTT and Hubble Deep Fields.

J/AJ/120/2269/  table1.dat table3.dat table4.dat  (58Kb)

Pinkney J., Burns J.O., Ledlow M.J., Gomez P.L., Hill J.M. Astron. J. 120, 2269 (2000) Substructure in clusters containing Wide-Angle-Tailed radio galaxies. I. New redshifts.

J/AJ/120/2331/  table1.dat  (41Kb)

Thompson D., Aftreth O., Soifer B.T. Astron. J. 120, 2331 (2000) B3 0003+387: AGN-marked large-scale structure at redshift 1.47 ?

J/AJ/120/2338/  table1.dat table2.dat  (28Kb)

Rines K., Geller M.J., Diaferio A., Mohr J.J., Wegner G.A. Astron. J. 120, 2338 (2000) The infall region of Abell 576: independent mass and light profiles.

J/AJ/120/2496/  table1.dat table4.dat  (3.3 Mb)

Dohm-Palmer R., Mateo M., Olszewski E., Morrison H., Harding P., Freeman K., Norris N. Astron. J., 120, 2496 (2000) Washington Stellar Photometry for the Spaghetti Survey

J/AJ/120/2638/  table2.dat table3.dat  (36Kb)

Horch E., Franz O.G., Ninkov Z. Astron. J. 120, 2638 (2000) CCD speckle observations of binary stars from the southern hemisphere. II. Measures from the Lowell-Tololo telescope during 1999.

J/AJ/120/2747/  table6.dat  (262Kb)

Casertano S., De Mello D.F., Dickinson M., Ferguson H.C., Fruchter A.S., Gonzalez-Lopezlira R.A., Heyer I., Hook R.N., Levay Z., Lucas R.A., Mack J., Makidon R.B., Mutchler M., Smith T.E., Stiavelli M., Wiggs M.S., Williams R.E. Astron. J. 120, 2747 (2000) WFPC2 observations of the Hubble Deep Field South.

J/AJ/121/31/  table1.dat table3.dat table4.dat table7.dat  (23Kb)

Fan X., Strauss M.A., Richards G.T., Newman J.A., Becker R.H., Schneider D.P., Gunn J.E., Davis M., White R.L., Lupton R.H., Anderson J.E.JR, Annis J., Bahcall N.A., Brunner R.J., Csabai I., Doi M., Fukugita M., Hennessy G.S., Hindsley R.B., Ivezic Z., Knapp G.R., Mckay T.A., Munn J.A., Pier J.R., Szalay A.S., York D.G. Astron. J., 121, 31-53 (2001) High-redshift quasars found in Sloan Digital Sky Survey commissioning data. III. A color-selected sample at i* 20 in the fall equatorial stripe.

J/AJ/121/66/  table1.dat table2.dat  (102Kb)

Salzer J.J., Gronwall C., Lipovetsky V.A., Kniazev A., Moody J.W., Boroson T.A., Thuan T.X., Izotov Y.I., Herrero J.L., Frattare L.M.

Astron. J. 121, 66 (2001) The KPNO International Spectroscopic Survey. II. Hα-selected survey list 1.

J/AJ/121/90/  table4.dat table5.dat  (20Kb)

Clements S.D., Carini M.T. Astron. J., 121, 90-96 (2001) Multiband microvariability observations of BL Lacertae during the outburst of 1997.

J/AJ/121/97/  table1.dat table5.dat table6.dat table7.dat table8.dat  (17Kb)

Murphy T.W.JR, Soifer B.T., Matthews K., Armus L., Kiger J.R. Astron. J., 121, 97-127 (2001) K-band spectroscopy of ultraluminous infrared galaxies: the 2 Jy sample.

J/AJ/121/148/  table3.dat  (12Kb)

Carrasco E.R., De Oliveira C.M., Infante L., Bolte M. Astron. J., 121, 148-168 (2001) The dwarf galaxy population of the Dorado group down to Mv = -11.

J/AJ/121/169/  table2.dat  (8Kb)

Fricke K.J., Izotov Y.I., Papaderos P., Guseva N.G., Thuan T.X. Astron. J., 121, 169-181 (2001) An imaging and spectroscopic study of the very metal-deficient blue compact dwarf galaxy Tol 1214-277.

J/AJ/121/182/  table2.dat  (5Kb)

Elmegreen D.M., Kaufman M., Elmegreen B.G., Brinks E., Struck C., Klaric M., Thomasson M. Astron. J., 121, 182-197 (2001) Hubble space telescope observations of dust and star-forming regions in the ocular galaxy IC 2163 and its spiral companion NGC 2207.

J/AJ/121/225/  table3.dat table4.dat table6.dat table8.dat  (10Kb)

Buta R., Ryder S.D., Madsen G.J., Wesson K., Crocker D.A., Combes F. Astron. J., 121, 225-243 (2001) Dynamics of ringed barred spiral galaxies. I. Surface photometry and kinematics of NGC 1433 and NGC 6300.

J/AJ/121/254/  table1.dat table2.dat  (13Kb)

Rucinski S.M., Maceroni C. Astron. J., 121, 254-266 (2001) Eclipsing binaries in the OGLE variable star catalogs. V. Long-period EB-type light curve systems in the Small Magellanic Cloud and the PLC-βrelation.

J/AJ/121/318/  table1.dat table2.dat  (318Kb)

Alves D.R., Bond H.E., Onken C. Astron. J. 121, 318 (2001) CCD photometry of the globular cluster NGC 5986 and its post-asymptotic giant branch and RR Lyrae stars.

J/AJ/121/327/  table3.dat  (18Kb)

Bragaglia A., Carretta E., Gratton R.G., Tosi M., Bonanno G., Bruno P., Cali A., Claudi R., Cosentino R., Desidera S., Farisato G., Rebeschini M., Scuderi S. Astron. J., 121, 327-336 (2001) Metal abundances of red clump stars in open clusters. I. NGC6819.

J/AJ/121/337/  table1.dat table2.dat  (11Kb)

Krishnamurthi A., Reynolds C.S., Linsky J.L., Martin E., Gagne M. Astron. J., 121, 337-346 (2001) Observations of the core of the Pleiades with the Chandra X-ray observatory.

J/AJ/121/362/  table1.dat  (8Kb)

Rudy R.J., Lynch D.K., Mazuk S., Puetter R.C., Dearborn D.S.P. Astron. J., 121, 362-370 (2001) The near-infrared spectrum of the planetary nebula IC 5117.

J/AJ/121/413/  table5.dat  (6Kb)

Arnal E.M. Astron. J., 121, 413-425 (2001) A high-resolution HI study of the interstellar medium local to HD193793.

J/AJ/121/768/  table2.dat table3.dat  (14Kb)

de Grijs R., O’Connell R.W., Gallagher III J.S. Astron. J. 121, 768 (2001) The fossil starburst in M82.

J/AJ/121/841/  table1.dat table4.dat  (7.0 Mb)

Piatek S., Pryor C., Armandroff T.E., Olszewski E.W. Astron. J. 121, 841 (2001) Stars of the Draco dwarf spheroidal galaxy beyond its measured tidal boundary.

J/AJ/121/861/  table1.dat table3.dat  (4.9 Mb)

Macri L.M., Stanek K.Z., Sasselov D.D., Krockenberger M., Kaluzny J. Astron. J. 121, 861 (2001) The DIRECT project: catalogs of stellar objects in nearby galaxies. I. The central part of M33.

J/AJ/121/870/ table1.dat table2.dat table3.dat table4.dat table5.dat table6.dat table7.dat table8.dat table9.dat table10.dat (4.2 Mb) Macri L.M., Stanek K.Z., Sasselov D.D., Krockenberger M., Kaluzny J. Astron. J. 121, 870 (2001) DIRECT distances to nearby galaxies using detached eclipsing binaries and Cepheids. VI. Variables in the central part of M33.

J/AJ/121/891/  table1.dat table3.dat  (36Kb)

Parker J.W., Zaritsky D., Stecher T.P., Harris J., Massey P. Astron. J. 121, 891 (2001) Ultraviolet and optical observations of OB associations and field stars in the southwest region of the Large Magellanic Cloud.

J/AJ/121/916/  table4a.dat table4b.dat table4c.dat table4d.dat  (3.0 Mb)

Testa V., Corsi C.E., Andreuzzi G., Iannicola G., Marconi G., Piersimoni A.M., Buonanno R. Astron. J., 121, 916-934 (2001) Horizontal-branch morphology and dense environments: Hubble Space Telescope observations of globular clusters NGC 2298, 5897, 6535, and 6626.

J/AJ/121/935/  table34.dat  (48Kb)

Siegel M.H., Majewski S.R., Cudworth K.M., Takamiya M. Astron. J., 121, 935-950 (2001) A cluster’s last stand: the death of Palomar 13.

J/AJ/121/974/  table3.dat table4.dat  (12Kb)

Gomez M., Kenyon S.J. Astron. J. 121, 974 (2001) A near-infrared imaging survey of the Chamaeleon I dark cloud.

J/AJ/121/1040/  table1.dat  (9Kb)

Preibisch T., Guenther E., Zinnecker H. Astron. J. 121, 1040 (2001) A large spectroscopic survey for young low-mass members of the upper Scorpius OB association.

J/AJ/121/1192/  table1.dat table2.dat  (11Kb)

McGary R.S., Brisken W.F., Fruchter A.S., Goss W.M., Thorsett S.E. Astron. J. 121, 1192 (2001) Proper-motion measurements with the VLA. I. Wide-field imaging and pulse-gating techniques.

J/AJ/121/1676/  table3.dat table4.dat  (45Kb)

Rebull L.M. Astron. J. 121, 1676 (2001) Rotation of young low-mass stars in the Orion Nebula Cluster flanking fields.

J/AJ/121/1840/  table4.dat  (11Kb)

Hall P.B., Sawicki M., Martini P., Finn R.A., Pritchet C.J., Osmer P.S., McCarthy D.W., Evans A.S., Lin H., Hartwick F.D.A. Astron. J. 121, 1840 (2001) Galaxies in the fields of z 1.5 radio-loud quasars.

J/AJ/121/1872/  table2.dat  (33Kb)

Helfand D.J., Stone R.P.S., Willman B., White R.L., Becker R.H., Price T., Gregg M.D., McMahon R.G. Astron. J. 121, 1872 (2001) Long-term optical variability of radio-selected quasars from the FIRST survey.

J/AJ/121/2020/  table1.dat table3.dat  (238Kb)

Bianchi L., Scuderi S., Massey P., Romaniello M. Astron. J. 121, 2020 (2001) The massive star content of NGC 6822: ground-based and Hubble Space Telescope photometry.

J/AJ/121/2032/ table2.dat table3.dat table4.dat table5.dat table6.dat table7.dat table8.dat table9.dat table10.dat table11.dat (5.6 Mb)

Mochejska B.J., Kaluzny J., Stanek K.Z., Sasselov D.D., Szentgyorgyi A.H. Astron. J. 121, 2032 (2001) DIRECT distances to nearby galaxies using detached eclipsing binaries and Cepheids. VII. Additional variables in the field M33A discovered with image subtraction.

J/AJ/121/2124/  table1.dat table2.dat table4.dat table5.dat table7.dat  (50Kb)

Dolan C.J., Mathieu R.D. Astron. J. 121, 2124 (2001) The spatial distribution of the λOrionis pre-main-sequence population.

J/AJ/121/2148/  notes.dat table1.dat  (38Kb)

Gray R.O., Napier M.G., Winkler L.I. Astron. J. 121, 2148 (2001) The physical basis of luminosity classification in the late A-, F-, and early G-type stars. I. Precise spectral types for 372 stars.

J/AJ/121/2159/  table1.dat  (16Kb)

Gray R.O., Graham P.W., Hoyt S.R. Astron. J. 121, 2159 (2001) The physical basis of luminosity classification in the late A-, F-, and early G-type stars. II. Basic parameters of program stars and the role of microturbulence.

J/AJ/121/2189/  table1.dat  (23Kb)

Oppenheimer B.R., Golimowski D.A., Kulkarni S.R., Matthews K., Nakajima T., Creech-Eakman M., Durrance S.T. Astron. J. 121, 2189 (2001) A coronagraphic survey for companions of stars within 8 parsecs.

J/AJ/121/2308/  table2.dat  (125Kb)

Richards G.T., Fan X., Schneider D.P., Vanden Berk D.E., Strauss M.A., York D.G., Anderson J.E.Jr, Anderson S.F., Annis J., Bahcall N.A., Bernardi M., Briggs J.W., Brinkmann J., Brunner R., Burles S., Carey L., Castander F.J., Connolly A.J., Crocker J.H., Csabai I., Doi M., Finkbeiner D., Friedman S.D., Frieman J.A., Fukugita M., Gunn J.E., Hindsley R.B., Ivezic Z., Kent S., Knapp G.R., Lamb D.Q., Leger R.F., Long D.C., Loveday J., Lupton R.H., McKay T.A., Meiksin A., Merrelli A., Munn J.A., Newberg H.J., Newcomb M., Nichol R.C., Owen R., Pier J.R., Pope A., Richmond M.W., Rockosi C.M., Schlegel D.J., Siegmund W.A., Smee S., Snir Y., Stoughton C., Stubbs C., Subbarao M., Szalay A.S., Szokoly G.P., Tremonti C., Uomoto A., Waddell P., Yanny B., Zheng W. Astron. J. 121, 2308 (2001) Colors of 2625 quasars at 0 z 5 measured in the Sloan Digital Sky Survey photometric system.

J/AJ/121/2331/  table1.dat table2.dat  (39Kb)

Castander F.J., Nichol R.C., Merrelli A., Burles S., Pope A., Connolly A.J., Uomoto A., Gunn J.E., Anderson J.E., Annis J., Bahcall N.A., Boroski W.N., Brinkmann J., Carey L., Crocker J.H., Csabai I., Doi M., Frieman J.A., Fukugita M., Friedman S.D., Hilton E.J., Hindsley R.B., Ivezi Z., Kent S., Lamb D.Q., Leger R.F., Long D.C., Loveday J., Lupton R.H., MacGillivray H., Meiksin A., Munn J.A., Newcomb M., Okamura S., Owen R., Pier J.R., Rockosi C.M., Schlegel D.J., Schneider D.P., Seigmund W., Smee S., Snir Y., Starkman L., Stoughton C., Szokoly G.P., Stubbs C., SubbaRao M., Szalay A., Thakar A.R., Tremonti C., Waddell P., Yanny B., York D.G. Astron. J. 121, 2331 (2001) The First Hour of Extra-galactic Data of the Sloan Digital Sky Survey Spectroscopic Commissioning: The Coma Cluster.

J/AJ/121/2381/  table2.dat  (24Kb)

Brown M.J.I., Webster R.L., Boyle B.J. Astron. J. 121, 2381 (2001) The evolution of radio galaxies at intermediate redshift.

J/AJ/121/2673/  table1.dat  (19Kb)

Kenyon S.J., Gomez M. Astron. J. 121, 2673 (2001) A 3μm survey of the Chamaeleon I dark cloud.

J/AJ/121/2815 See I/273A/

J/AJ/121/2843/  table1.dat table2.dat  (558Kb)

Barkhouse W.A., Hall P.B. Astron. J. 121, 2843 (2001) Quasars in the 2MASS second incremental data release.

J/AJ/121/2895/  table1.dat table6a.dat table6b.dat table7a.dat table7b.dat table8.dat  (54Kb)

Cohen J.G. Astron. J. 121, 2895 (2001) Caltech Faint Galaxy Redshift Survey. XIII. Spectral energy distributions for galaxies in the region of the Hubble Deep Field North.

J/AJ/121/3048/  table2.dat table3.dat table4.dat table5.dat table6.dat  (156Kb)

Maoz D., Barth A.J., Ho L.C., Sternberg A., Filippenko A.V. Astron. J. 121, 3048 (2001) An ultraviolet through infrared look at star formation and super star clusters in two circumnuclear starburst rings.

J/AJ/121/3114/  table1.dat table2.dat  (13Kb)

Paltrinieri B., Ferraro F.R., Paresce F., de Marchi G. Astron. J. 121, 3114 (2001) VLT observations of the peculiar globular cluster NGC 6712. III. The evolved stellar population.

J/AJ/121/3160/  table4.dat table7.dat  (123Kb)

Carpenter J.M., Hillenbrand L.A., Skrutskie M.F. Astron. J. 121, 3160 (2001) Near-infrared photometric variability of stars toward the Orion A molecular cloud.

J/AJ/121/3224/  refs.dat table1.dat table2.dat table3.dat table4.dat table5.dat table6.dat  (50Kb)

Mason B.D., Hartkopf W.I., Holdenried E.R., Rafferty T.J. Astron. J. 121, 3224 (2001) Speckle interferometry of new and problem Hipparcos binaries. II. Observations obtained in 1998-1999 from McDonald observatory.

J/AJ/122/38/  table3.dat  (11Kb)

Trujillo I., Aguerri J.A.L., Gutierrez C.M., Cepa J. Astron. J. 122, 38 (2001) Quantitative morphology of the intermediate-redshift galaxy cluster Abell 2443 from ground-based imaging: evidence for a galaxy concentration index correlation with cluster density.

J/AJ/122/248/  table1.dat table2.dat table4.dat  (1.1 Mb)

Keller S.C., Grebel E.K., Miller G.J., Yoss K.M. Astron. J. 122, 248 (2001) UBVI and Hαphotometry of the h and χPersei cluster.

J/AJ/122/503/  table1.dat table2.dat  (19Kb)

Anderson S.F., Fan X., Richards G.T., Schneider D.P., Strauss M.A., Vanden Berk D.E., Gunn J.E., Knapp G.R., Schlegel D., Voges W., Yanny B., Bahcall N.A., Bernardi M., Brinkmann J., Brunner R., Csabai I., Doi M., Fukugita M., Hennessy G.S., Ivezic Z., Kunszt P.Z., Lamb D.Q., Loveday J., Lupton R.H., McKay T.A., Munn J.A., Nichol R.C., Szokoly G.P., York D.G. Astron. J. 122, 503 (2001) High-redshift quasars found in Sloan Digital Sky Survey commissioning data. VI. Sloan Digital Sky Survey spectrograph observations.

J/AJ/122/518/  table1.dat table2.dat table3.dat table4.dat table5.dat  (38Kb)

Hewett P.C., Foltz C.B., Chaffee F.H. Astron. J. 122, 518 (2001) The Large Bright Quasar Survey. VII. The LBQS and FIRST surveys.

J/AJ/122/598/  notes.dat table1.dat table2.dat  (16Kb)

Dawson S., Stern D., Bunker A.J., Spinrad H., Dey A. Astron. J. 122, 598 (2001) Serendipitously detected galaxies in the Hubble Deep Field.

J/AJ/122/611/  table1.dat  (15Kb)

van den Bergh S., Cohen J.G., Crabbe C. Astron. J. 122, 611 (2001) Caltech Faint Galaxy Redshift Survey. XV. Classifications of galaxies with 0.2 Z 1.1 in the Hubble Deep Field North and its flanking fields.

J/AJ/122/621/  table1.dat  (12Kb)

van den Bergh S. Astron. J. 122, 621 (2001) The evolution of galaxy morphology for galaxies in the Canada-France Redshift Survey.

J/AJ/122/714/  table3.dat  (102Kb)

Brown W.R., Geller M.J., Fabricant D.G., Kurtz M.J. Astron. J. 122, 714 (2001)

V- and R-band galaxy luminosity functions and low surface brightness galaxies in the Century Survey.

J/AJ/122/866/  table1.dat table2.dat  (21Kb)

Preibisch T., Zinnecker H. Astron. J. 122, 866 (2001) Deep Chandra X-ray observatory imaging study of the very young stellar cluster IC 348.

J/AJ/122/1125/  table2.dat table3.dat table4.dat  (63Kb)

Postman M., Lubin L.M., Oke J.B. Astron. J. 122, 1125 (2001) A study of nine high-redshift clusters of galaxies. IV. Photometry and spectra of clusters 1324+3011 and 1604+4321.

J/AJ/122/1289/  table1.dat table2.dat  (43Kb)

Tustin A.W., Geller M.J., Kenyon S.J., Diaferio A. Astron. J. 122, 1289 (2001) H-band and spectroscopic properties of A1644.

J/AJ/122/1330/  table2.dat table3.dat  (18Kb)

Chandar R., Ford H.C., Tsvetanov Z. Astron. J. 122, 1330 (2001) Compact star clusters in M81. I. Data from Hubble Space Telescope Wide Field Planetary Camera 2 imaging.

J/AJ/122/1383/  table1a.dat table1b.dat table1c.dat table1d.dat table1e.dat  (2.3 Mb)

Mochejska B.J., Kaluzny J., Stanek K.Z., Sasselov D.D. Astron. J. 122, 1383 (2001) The DIRECT project: catalogs of stellar objects in nearby galaxies. II. Eastern arm and NGC 206 in M31.

J/AJ/122/1486/  ccd.dat spm.dat table1.dat table3.dat  (1.8 Mb)

Platais I., Kozhurina-Platais V., Barnes S., Girard T.M., Demarque P., van Altena W.F., Deliyannis C.P., Horch E. Astron. J. 122, 1486 (2001) WIYN open cluster study. VII. NGC 2451A and the Hipparcos distance scale.

J/AJ/122/1586/  table1.dat table2.dat  (94Kb)

Mason B.D., Hartkopf W.I., Wycoff G.L., Holdenried E.R., Platais I., Rafferty T.J., Hall D.M., Hennessy G.S., Urban S.E., Pohl B.L. Astron. J. 122, 1586 (2001) Speckle interferometry at the US naval observatory. VII.

J/AJ/122/1718/  table1.dat table2.dat table3.dat  (41Kb)

Yuan Q., Zhou X., Chen J., Jiang Z., Ma J., Wu H., Xue S., Zhu J. Astron. J. 122, 1718 (2001) Multicolor photometry of the galaxies in the central region of Abell 2634.

J/AJ/122/1844/  table2.dat table3.dat table4.dat  (140Kb)

Egan M.P., van Dyk S.D., Price S.D. Astron. J. 122, 1844 (2001) MSX, 2MASS, and the Large Magellanic Cloud: a combined near- and mid-infrared view.

J/AJ/122/2055/  bands.dat table1.dat table2.dat table8.dat table9.dat  (91Kb)

Gonzalez-Perez J.N., Kidger M.R., Martin-Luis F. Astron. J. 122, 2055 (2001) Optical and near-infrared calibration of AGN field stars: an all-sky network of faint stars calibrated on the Landolt system.

J/AJ/122/2205/  table1.dat table3.dat  (37Kb)

Rudnick G., Franx M., Rix H.-W., Moorwood A., Kuijken K., van Starkenburg L., van der Werf P., Rottgering H., van Dokkum P., Labbe I. Astron. J. 122, 2205 (2001) A K-band-selected photometric redshift catalog in the Hubble Deep Field South: sampling the rest-frame V band to z=3.

J/AJ/122/2222/  tablea1.dat tableb1.dat tableb2.dat  (57Kb)

Einasto M., Einasto J., Tago E., Muller V., Andernach H. Astron. J. 122, 2222 (2001) Optical and X-ray clusters as tracers of the supercluster-void network. I. Superclusters of Abell and X-ray clusters.

J/AJ/122/2458/  table1.dat table2.dat table3.dat table4.dat  (32Kb)

Barmby P., Huchra J.P.

Astron. J. 122, 2458 (2001) M31 globular clusters in the Hubble Space Telescope archive. I. Cluster detection and completeness.

J/AJ/122/2477/ table1.dat table2.dat table3.dat table4.dat table5.dat table6.dat table7.dat table8.dat table9.dat table10.dat (4.7 Mb) Mochejska B.J., Kaluzny J., Stanek K.Z., Sasselov D.D., Szentgyorgyi A.H. Astron. J. 122, 2477 (2001) DIRECT distances to nearby galaxies using detached eclipsing binaries and Cepheids. VIII. Additional variables in the field M33B discovered with image subtraction.

J/AJ/122/2569/  table1.dat table2.dat table3.dat  (1.6 Mb)

Bellazzini M., Pecci F.F., Ferraro F.R., Galleti S., Catelan M., Landsman W.B. Astron. J. 122, 2569 (2001) Age as the second parameter in NGC 288/NGC 362? I. Turnoff ages: a purely differential comparison.

J/AJ/122/2810/  table3.dat table6.dat  (71Kb)

Brandt W.N., Alexander D.M., Hornschemeier A.E., Garmire G.P., Schneider D.P., Barger A.J., Bauer F.E., Broos P.S., Cowie L.L., Townsley L.K., Burrows D.N., Chartas G., Feigelson E.D., Griffiths R.E., Nousek J.A., Sargent W.L.W. Astron. J. 122, 2810 (2001) The Chandra Deep Field North survey. V. 1 Ms source catalogs.

J/AJ/122/2893/  table2.dat table3.dat  (125Kb)

Wegner G., Thorstensen J.R., Kurtz M.J., Brown W.R., Fabricant D.G., Geller M.J., Huchra J.P., Marzke R.O., Sakai S. Astron. J. 122, 2893 (2001) Redshifts for 2410 galaxies in the Century Survey region.

J/AJ/122/3046/  table2.dat table3.dat  (27Kb)

Harris J., Calzetti D., Gallagher III J.S., Conselice C.J., Smith D.A. Astron. J. 122, 3046 (2001) Young clusters in the nuclear starburst of M83.

J/AJ/122/3155/  table3.dat table4.dat table5.dat  (22Kb)

Higgs L.A., Kerton C.R., Knee L.B.G. Astron. J. 122, 3155 (2001) A peculiar intermediate-velocity hydrogen cloud near the galactic plane.

J/AJ/122/3219/  table1.dat table2.dat  (218Kb)

Rey S.-C., Yoon S.-J., Lee Y.-W., Chaboyer B., Sarajedini A. Astron. J. 122, 3219 (2001) CCD photometry of the classic second-parameter globular clusters M3 and M13.

J/AJ/122/3482 See II/235/

J/AJ/123/51/  tablea1.dat  (16Kb)

Einasto M., Einasto J., Tago E., Andernach H., Dalton G.B., Muller V. Astron. J. 123, 51 (2002) Optical and X-ray clusters as tracers of the supercluster-void network. III. Distribution of Abell and APM clusters.

J/AJ/123/100/  table1.dat table2.dat  (109Kb)

Koranyi D.M., Geller M.J. Astron. J. 123, 100 (2002) Kinematics of AWM and MKW poor clusters.

J/AJ/123/269/  table2.dat table3.dat  (8Kb)

Jacoby G.H., De Marco O. Astron. J., 123, 269-278 (2002) A survey for very faint planetary nebulae in the SMC. I. Identification, confirmation, and preliminary analysis.

J/AJ/123/279/ phot.dat table1.dat v1 i.dat v1 v.dat v2 i.dat v2 v.dat v3 i.dat v3 v.dat v4 i.dat v4 v.dat v5 i.dat v5 v.dat v6 i.dat v6 v.dat v7 i.dat v7 v.dat v8 i.dat v8 v.dat v9 i.dat v9 v.dat v10 i.dat v10 v.dat v11 i.dat v11 v.dat v12 i.dat v12 v.dat v13 i.dat v13 v.dat v14 i.dat v14 v.dat (1.6 Mb) von Braun K., Mateo M. Astron. J., 123, 279 (2002) Variable stars in the field of the globular cluster NGC 3201

J/AJ/123/290/  table1.dat table4.dat  (184Kb)

Sung H., Bessell M.S., Lee B.-W., Lee S.-G. Astron. J. 123, 290 (2002)

The open cluster NGC 2516. I. Optical photometry.

J/AJ/123/304/  table1.dat table2.dat table3.dat table4.dat  (80Kb)

Herbig G.H., Dahm S.E. Astron. J. 123, 304 (2002) The young cluster IC 5146.

J/AJ/123/387/  fields.dat phot.dat  (23.0 Mb)

Dolan C.J., Mathieu R.D. Astron. J. 123, 387 (2002) A photometric study of the young stellar population throughout the λOrionis star-forming region.

J/AJ/123/567/  table1.dat table3.dat table5.dat  (1Kb)

Schneider D.P., Richards G.T., Fan X., Hall P.B., Strauss M.A., Vanden Berk D.E., Gunn J.E., Newberg H.J., Reichard T.A., Stoughton C., Voges W., Yanny B., Anderson S.F., Annis J., Bahcall N.A., Bauer A., Bernardi M., Blanton M.R., Boroski W.N., Brinkmann J., Briggs J.W., Brunner R., Burles S., Carey L., Castander F.J., Connolly A.J., Csabai I., Doi M., Friedman S., Frieman J.A., Fukugita M., Heckman T.M., Hennessy G.S., Hindsley R.B., Hogg D.W., Ivezic Z., Kent S., Knapp G.R., Kunzst P.Z., Lamb D.Q., Leger R.F., Long D.C., Loveday J., Lupton R.H., Margon B., Meiksin A., Merelli A., Munn J.A., Newcomb M., Nichol R.C., Owen R., Pier J.R., Pope A., Rockosi C.M., Saxe D.H., Schlegel D., Siegmund W.A., Smee S., Snir Y., Subbarao M., Szalay A.S., Thakar A.R., Uomoto A., Waddell P., York D.G. Astron. J. 123, 567 (2002) The Sloan Digital Sky Survey quasar catalog. I. Early data release.

J/AJ/123/578/  table1.dat table2.dat  (11Kb)

Chen Y., He X.-T., Wu J.-H., Li Q.-K., Green R.F., Voges W. Astron. J. 123, 578 (2002) The Multiwavelength Quasar Survey. II. Quasars in the Coma cluster.

J/AJ/123/637/  table5.dat table6.dat table7.dat  (39Kb)

De Breuck C., van Breugel W., Stanford S.A., Rottgering H., Miley G., Stern D. Astron. J. 123, 637 (2002) Optical and near-infrared imaging of ultra-steep-spectrum radio sources: the K-z diagram of radio-selected and optically selected galaxies.

J/AJ/123/789/  table1.dat table2.dat table3.dat table4.dat  (211Kb)

Pietrzynski G., Gieren W., Fouque P., Pont F. Astron. J. 123, 789 (2002) The ARAUCARIA project: discovery of Cepheid variables in NGC 300 from a wide-field imaging survey.

J/AJ/123/832/  table2.dat table3.dat  (78Kb)

Letarte B., Demers S., Battinelli P., Kunkel W.E. Astron. J. 123, 832 (2002) The extent of NGC 6822 revealed by its C star population.

J/AJ/123/840/  table1.dat table2.dat table3.dat table4.dat  (62Kb)

Bersier D., Wood P.R. Astron. J. 123, 840 (2002) Variable stars in the Fornax dwarf galaxy.

J/AJ/123/855/  table1.dat  (574.0 Mb)

Zaritsky D., Harris J., Thompson I.B., Grebel E.K., Massey P. Astron. J. 123, 855 (2002) The Magellanic Clouds Photometric Survey: the Small Magellanic Cloud stellar catalog and extinction map.

J/AJ/123/873/  table1.dat table3.dat  (263Kb)

Putman M.E., de Heij V., Staveley-Smith L., Braun R., Freeman K.C., Gibson B.K., Burton W.B., Barnes D.G., Banks G.D., Bhathal R., de Blok W.J.G., Boyce P.J., Disney M.J., Drinkwater M.J., Ekers R.D., Henning P.A., Jerjen H., Kilborn V.A., Knezek P.M., Koribalski B., Malin D.F., Marquarding M., Minchin R.F., Mould J.R., Oosterloo T., Price R.M., Ryder S.D., Sadler E.M., Stewart I., Stootman F., Webster R.L., Wright A.E. Astron. J. 123, 873 (2002) HIPASS high-velocity clouds: properties of the compact and extended populations.

J/AJ/123/1023/  table1.dat table2.dat  (25Kb)

Seymour D.M., Mason B.D., Hartkopf W.I., Wycoff G.L. Astron. J. 123, 1023 (2002) Binary star orbits. II. Preliminary first orbits for 117 systems.

J/AJ/123/1200/  notes.dat refs.dat table2.dat table3.dat  (26Kb)

Caretta C.A., Maia M.A.G., Kawasaki W., Willmer C.N.A.

Astron. J. 123, 1200 (2002) The Aquarius superclusters. I. Identification of clusters and superclusters.

J/AJ/123/1216/  table1.dat  (28Kb)

Rose J.A., Gaba A.E., Christiansen W.A., Davis D.S., Caldwell N., Hunstead R.W., Johnston-Hollitt M. Astron. J. 123, 1216 (2002) Multiple merging events in the double cluster A3128/A3125.

J/AJ/123/1292/  table2.dat table3.dat  (39Kb)

Salzer J.J., Gronwall C., Sarajedini V.L., Lipovetsky V.A., Kniazev A., Moody J.W., Boroson T.A., Thuan T.X., Izotov Y.I., Herrero J.L., Frattare L.M. Astron. J. 123, 1292 (2002) The KPNO International Spectroscopic Survey. III. [O III]-selected survey list.

J/AJ/123/1411/  table2.dat table4.dat table5.dat  (19Kb)

Benedict G.F., Howell D.A., Jorgensen I., Kenney J.D.P., Smith B.J. Astron. J. 123, 1411 (2002) NGC 4314. IV. Photometry of star clusters with the Hubble space telescope: history of star formation in the vicinity of a nuclear ring.

J/AJ/123/1528/  table1.dat table2.dat table3.dat  (119Kb)

Rebull L.M., Makidon R.B., Strom S.E., Hillenbrand L.A., Birmingham A., Patten B.M., Jones B.F., Yagi H., Adams M.T. Astron. J. 123, 1528 (2002) Circumstellar disk candidates identified in NGC 2264.

J/AJ/123/1570/ table1.dat table2.dat table4a.dat table4b.dat table5a.dat table6.dat table7.dat (55Kb) Patience J., Ghez A.M., Reid I.N., Matthews K. Astron. J. 123, 1570 (2002) A high angular resolution multiplicity survey of the open clusters αPersei and Praesepe.

J/AJ/123/1784/ 6sigma.dat 7sigma.dat 8sigma.dat 9sigma.dat 10sigma.dat booapmm.dat booapms.dat boowenss.dat catalog.dat table2.dat (1.8 Mb) de Vries W.H., Morganti R., Rottgering H.J.A., Vermeulen R., van Breugel W., Rengelink R., Jarvis M.J. Astron. J. 123, 1784 (2002) Deep Westerbork 1.4 GHz imaging of the Bootes field.

J/AJ/123/1807/  v4-7526.dat v4-94125.dat  (255Kb)

Goto T., Sekiguchi M., Nichol R.C., Bahcall N.A., Kim R.S.J., Annis J., Ivezic Z., Brinkmann J., Hennessy G.S., Szokoly G.P., Tucker D.L. Astron. J. 123, 1807 (2002) The cut-and-enhance method: selecting clusters of galaxies from the Sloan Digital Sky Survey commissioning data.

J/AJ/123/2070/  table2.dat table3.dat  (67Kb)

Tedesco E.F., Desert F.-X. Astron. J. 123, 2070 (2002) The Infrared Space Observatory deep asteroid search.

J/AJ/123/2206/  table4.dat table5.dat table6.dat  (917Kb)

Prochaska J.X., Gawiser E., Wolfe A.M., Quirrenbach A., Lanzetta K.M., Chen H.-W., Cooke J., Yahata N. Astron. J. 123, 2206 (2002) Galaxies associated with z 4 damped Ly αsystems. I. Imaging and photometric selection.

J/AJ/123/2223/  table1.dat table2.dat table3.dat  (36Kb)

Stern D., Tozzi P., Stanford S.A., Rosati P., Holden B., Eisenhardt P., Elston R., Wu K.L., Connolly A., Spinrad H., Dawson S., Dey A., Chaffee F.H. Astron. J. 123, 2223 (2002) SPICES II: optical and near-infrared identifications of faint X-ray sources from deep CHANDRA observations of Lynx.

J/AJ/123/2261/  table3.dat table10.dat  (28Kb)

Berrington R.C., Lugger P.M., Cohn H.N. Astron. J. 123, 2261 (2002) The dynamics of the merging galaxy cluster system A2256: evidence for a new subcluster.

J/AJ/123/2402/  table1.dat table2.dat table4.dat  (21Kb)

Fomalont E.B., Kellermann K.I., Partridge R.B., Windhorst R.A., Richards E.A. Astron. J. 123, 2402 (2002) The microJansky sky at 8.4 GHz.

J/AJ/123/2490/  table3.dat  (29Kb)

Perrett K.M., Bridges T.J., Hanes D.A., Irwin M.J., Brodie J.P., Carter D., Huchra J.P., Watson F.G. Astron. J. 123, 2490 (2002) The kinematics and metallicity of the M31 globular cluster system.

J/AJ/123/2511/  table2.dat table3.dat  (1.4 Mb)

Piatek S., Pryor C., Armandroff T.E., Olszewski E.W. Astron. J. 123, 2511 (2002) Structure of the Draco dwarf spheroidal galaxy.

J/AJ/123/2597/  table4.dat table5.dat table7.dat  (33Kb)

Ogura K., Sugitani K., Pickles A. Astron. J. 123, 2597 (2002) Hαemission stars and Herbig-Haro objects in the vicinity of bright-rimmed clouds.

J/AJ/123/2772/  atca/* doublede.dat infrared.dat lbv.dat tablevla.dat vla/*  (9.4 Mb)

Sevenster M.N. Astron. J. 123, 2772 (2002) OH-selected AGB and post-AGB objects. I. Infrared and maser properties.

J/AJ/123/2806/  ref3.dat table1.dat table2.dat table3.dat table4.dat table6.dat table7.dat  (126Kb)

Reid I.N., Cruz K.L. Astron. J. 123, 2806 (2002) Meeting the cool neighbors. I. Nearby stars in the NLTT catalogue: defining the sample.

J/AJ/123/2822/  table1.dat table3.dat  (45Kb)

Reid I.N., Kilkenny D., Cruz K.L. Astron. J. 123, 2822 (2002) Meeting the cool neighbors. II. Photometry of southern NLTT stars.

J/AJ/123/2828/  table1.dat table2.dat table3.dat table5.dat table6.dat  (42Kb)

Cruz K.L., Reid I.N. Astron. J. 123, 2828 (2002) Meeting the cool neighbors. III. Spectroscopy of northern NLTT stars.

J/AJ/123/2976/  table1.dat table2.dat table3.dat table4.dat  (408Kb)

Ramella M., Geller M.J., Pisani A., da Costa L.N. Astron. J. 123, 2976 (2002) The UZC-SSRS2 group catalog.

J/AJ/123/3018/  refs.dat table1.dat table2.dat table3.dat  (44Kb)

Miller N.A., Ledlow M.J., Owen F.N., Hill J.M. Astron. J. 123, 3018 (2002) Redshifts for a sample of radio-selected poor clusters.

J/AJ/123/3154/  table2.dat table3.dat  (100Kb)

Dolphin A.E., Saha A., Claver J., Skillman E.D., Cole A.A., Gallagher J.S., Tolstoy E., Dohm-Palmer R.C., Mateo M. Astron. J. 123, 3154 (2002) Variable stars in Leo A: RR Lyrae stars, short-period Cepheids, and implications for stellar content.

J/AJ/123/3216/  table2.dat table3.dat table5.dat table6.dat  (553Kb)

Sharpee B., Stark M., Pritzl B., Smith H., Silbermann N., Wilhelm R., Walker A. Astron. J. 123, 3216 (2002) BV photometry of variable stars in the northeast arm of the Small Magellanic Cloud.

J/AJ/123/3356/  table2.dat table4.dat table6.dat  (148Kb)

Gizis J.E., Reid I.N., Hawley S.L. Astron. J. 123, 3356 (2002) The Palomar/MSU nearby star spectroscopic survey. III. Chromospheric activity, M dwarf ages and the local star formation history.

J/AJ/123/3409/  ml/* sp/* spectra.dat table1.dat table3.dat table4.dat  (4.1 Mb)

Hawley S.L., Covey K.R., Knapp G.R., Golimowski D.A., Fan X., Anderson S.F., Gunn J.E., Harris H.C., Ivezic Z., Long G.M., Lupton R.H., McGehee P.M., Narayanan V., Peng E., Schlegel D., Schneider D.P., Spahn E.Y., Strauss M.A., Szkody P., Tsvetanov Z., Walkowicz L.M., Brinkmann J., Harvanek M., Hennessy G.S., Kleinman S.J., Krzesinski J., Long D., Neilsen E.H., Newman P.R., Nitta A., Snedden S.A., York D.G. Astron. J. 123, 3409 (2002) Characterization of M, L, and T dwarfs in the Sloan Digital Sky Survey.

J/AJ/123/3460/  lcurve/* table1.dat table2.dat table3.dat  (383Kb)

Mochejska B.J., Stanek K.Z., Sasselov D.D., Szentgyorgyi A.H. Astron. J. 123, 3460 (2002) Planets in stellar clusters extensive search. I. Discovery of 47 low-amplitude variables in the metal-rich cluster NGC 6791 with millimagnitude image subtraction photometry.

J/AJ/124/100/ notes.dat refs.dat table1.dat table2.dat table3.dat table4.dat table5.dat table6.dat table7.dat table8.dat tnotes.dat (74Kb) Darling J., Giovanelli R. Astron. J. 124, 100 (2002) A search for OH megamasers at z 0.1. III. The complete survey.

J/AJ/124/234/  table1.dat table3.dat  (24Kb)

Reitzel D.B., Guhathakurta P. Astron. J. 124, 234 (2002) Metallicity and kinematics of M31’s outer stellar halo from a Keck spectroscopic survey.

J/AJ/124/404/  table1.dat  (14Kb)

Preibisch T., Brown A.G.A., Bridges T., Guenther E., Zinnecker H. Astron. J. 124, 404 (2002) Exploring the full stellar population of the Upper Scorpius OB association.

J/AJ/124/445/  table1.dat table2.dat  (20Kb)

Doi T., O’dell C.R., Hartigan P. Astron. J., 124, 445-463 (2002) Internal velocities in the Orion nebula: large proper-motion features.

J/AJ/124/601/  ngc188.dat plates.dat  (130Kb)

Platais I., Kozhurina-Platais V., Girard T.M., van Altena W.F., Klemola A.R., Stauffer J.R., Armandroff T.E., Mighell K.J., Dell’Antonio I., Falco E.E., Sarajedini A. Astron. J. 124, 601 (2002) WIYN Open Cluster Study. VIII. The geometry and stability of the NOAO CCD Mosaic Imager

J/AJ/124/690/  table1.dat table2.dat  (67Kb)

Kilborn V.A., Webster R.L., Staveley-Smith L., Marquarding M., Banks G.D., Barnes D.G., Bhathal R., de Blok W.J.G., Boyce P.J., Disney M.J., Drinkwater M.J., Ekers R.D., Freeman K.C., Gibson B.K., Henning P.A., Howlett M., Jerjen H., Knezek P.M., Koribalski B., Malin D.F., Minchin R.F., Mould J.R., O’Brien J.C., Oosterloo T., Ortiz P.F., Price R.M., Putman M.E., Ryder S.D., Sadler E.M., Stewart I.M., Stootman F., Wong O.I., Wright A.E. Astron. J. 124, 690 (2002) A catalog of H I-selected galaxies from the south celestial cap region of sky.

J/AJ/124/782/  table1.dat table2.dat table3.dat  (22Kb)

van den Bergh S. Astron. J. 124, 782 (2002) Bar galaxies and their environments.

J/AJ/124/915/  table1.dat  (125Kb)

Bellazzini M., Ferraro F.R., Ibata R. Astron. J. 124, 915 (2002) The stellar population of NGC 5634: a globular cluster in the Sagittarius dSph stream?

J/AJ/124/1001/  table3.dat table4.dat table5.dat table6.dat  (74Kb)

Carpenter J.M., Hillenbrand L.A., Skrutskie M.F., Meyer M.R. Astron. J. 124, 1001 (2002) Near-infrared photometric variability of stars toward the Chamaeleon I molecular cloud.

J/AJ/124/1118/  table1.dat  (11Kb)

Silvestri N.M., Oswalt T.D., Hawley S.L. Astron. J. 124, 1118 (2002) Wide binary systems and the nature of high-velocity white dwarfs.

J/AJ/124/1144/  table1.dat table2.dat table3.dat table4.dat table5.dat  (1.9 Mb)

Latham D.W., Stefanik R.P., Torres G., Davis R.J., Mazeh T., Carney B.W., Laird J.B., Morse J.A. Astron. J. 124, 1144 (2002) A survey of proper-motion stars. XVI. Orbital solutions for 171 single-lined spectroscopic binaries.

J/AJ/124/1190/  table1.dat table3.dat table4.dat  (24Kb)

Lepine S., Shara M.M., Rich R.M. Astron. J. 124, 1190 (2002) New high proper motion stars from the Digitized Sky Survey. I. Northern stars with 0.5”/yr μ2.0”/yr at low galactic latitudes.

J/AJ/124/1266/  table3.dat table5.dat table6.dat  (53Kb)

Rines K., Geller M.J., Diaferio A., Mahdavi A., Mohr J.J., Wegner G. Astron. J. 124, 1266 (2002) Mass profile of the infall region of the Abell 2199 supercluster.

J/AJ/124/1283/  table1.dat table2.dat  (43Kb)

Stepanian J.A., Chavushyan V.H., Carrasco L., Valdes J.R., Mujica R.M., Tovmassian H.M., Ayvazyan V.T. Astron. J. 124, 1283 (2002) Spectral observations of faint Markarian galaxies of the Second Byurakan Survey. III.

J/AJ/124/1418/  notes.dat table2.dat table3.dat  (17Kb)

Whitmore B.C., Zhang Q. Astron. J. 124, 1418 (2002) What fraction of the young clusters in the Antennae galaxies are ”missing”?

J/AJ/124/1464/  table1.dat table2.dat table3.dat table4.dat  (108Kb)

Pritzl B.J., Armandroff T.E., Jacoby G.H., da Costa G.S. Astron. J. 124, 1464 (2002) The dwarf spheroidal companions to M31: variable stars in Andromeda VI.

J/AJ/124/1486/  m4ubv.dat m4wd.dat  (1.2 Mb)

Mochejska B.J., Kaluzny J., Thompson I., Pych W. Astron. J. 124, 1486 (2002) Clusters ages experiment: hot subdwarfs and luminous white dwarf candidates in the field of the globular cluster M4.

J/AJ/124/1570/  table1.dat table4.dat table5.dat table6.dat  (549Kb)

Adams J.D., Stauffer J.R., Skrutskie M.F., Monet D.G., Portegies Zwart S.F., Janes K.A., Beichman C.A. Astron. J. 124, 1570 (2002) Structure of the Praesepe star cluster.

J/AJ/124/1651/  table1.dat table2.dat  (11Kb)

Margon B., Anderson S.F., Harris H.C., Strauss M.A., Knapp G.R., Fan X., Schneider D.P., Vanden Berk D.E., Schlegel D.J., Deutsch E.W., Ivezic Z., Hall P.B., Williams B.F., Davidsen A.F., Brinkmann J., Csabai I., Hayes J.J.E., Hennessy G., Kinney E.K., Kleinman S.J., Lamb D.Q., Long D., Neilsen E.H., Nichol R., Nitta A., Snedden S.A., York D.G. Astron. J. 124, 1651 (2002) Faint high-latitude carbon stars discovered by the Sloan Digital Sky Survey: methods and initial results.

J/AJ/124/1670/  notes.dat table1.dat table2.dat table3.dat table4.dat table5.dat table6.dat  (67Kb)

Mamajek E.E., Meyer M.R., Liebert J. Astron. J. 124, 1670 (2002) Post-T Tauri Stars in the Nearest OB Association

J/AJ/124/1918/  notes.dat refs.dat table1.dat table2.dat  (135Kb)

Miller C.J., Krughoff K.S., Batuski D.J., Hill J.M. Astron. J. 124, 1918 (2002) The MX northern Abell cluster survey. II: the Abell/ACO cluster redshifts and spatial analyses.

J/AJ/124/1954/  notes.dat table2.dat table3.dat  (37Kb)

Ryan-Weber E., Koribalski B.S., Staveley-Smith L., Jerjen H., Kraan-Korteweg R.C., Ryder S.D., Barnes D.G., de Blok W.J.G., Kilborn V.A., Bhathal R., Boyce P.J., Disney M.J., Drinkwater M.J., Ekers R.D., Freeman K.C., Gibson B.K., Green A.J., Haynes R.F., Henning P.A., Juraszek S., Kesteven M.J., Knezek P.M., Mader S., Marquarding M., Meyer M., Minchin R.F., Mould J.R., O’Brien J., Oosterloo T., Price R.M., Putman M.E., Sadler E.M., Schroder A., Stewart I.M., Stootman F., Waugh M., Webster R.L., Wright A.E., Zwaan M.A. Astron. J. 124, 1954 (2002) The 1000 brightest HIPASS galaxies: newly catalogued galaxies.

J/AJ/124/2039/  table1.dat  (75Kb)

Keller S.C., Bessell M.S., Cook K.H., Geha M., Syphers D. Astron. J. 124, 2039 (2002) Blue variable stars from the MACHO database. I. Photometry and spectroscopy of the Large Magellanic Cloud sample.

J/AJ/124/2164/  table2.dat table3.dat table4.dat  (43Kb)

Andersson B.-G., Idzi R., Uomoto A., Wannier P.G., Chen B., Jorgensen A.M. Astron. J. 124, 2164 (2002) A spectroscopic and photometric survey of stars in the field of L1457: a new distance determination.

J/AJ/124/2254/  table1.dat table2.dat table3.dat  (96Kb)

Mason B.D., Hartkopf W.I., Urban S.E., Wycoff G.L., Pascu D., Hall D.M., Hennessy G.S., Platais I., Rafferty T.J., Holdenried E.R. Astron. J. 124, 2254 (2002) Speckle interferometry at the US naval observatory. VIII.

J/AJ/124/2351/  table1.dat table2.dat  (19Kb)

Bauer F.E., Alexander D.M., Brandt W.N., Hornschemeier A.E., Vignali C., Garmire G.P., Schneider D.P. Astron. J., 124, 2351-2363 (2002) The Chandra Deep Field North survey. XII. The link between faint X-ray and radio source populations.

J/AJ/124/2453/  table3.dat table4.dat table5.dat table6.dat table8.dat  (172Kb)

Miller N.A., Owen F.N. Astron. J. 124, 2453 (2002) Evolution of star-forming and active galaxies in nearby clusters.

J/AJ/124/2955/  table1.dat table2.dat  (6.1 Mb)

Grazian A., Omizzolo A., Corbally C., Cristiani S., Haehnelt M.G., Vanzella E. Astron. J. 124, 2955 (2002) The Asiago-ESO/RASS QSO survey. II. The southern sample.

J/AJ/124/2990/  table4.dat table5.dat  (42Kb)

Kraemer K.E., Price S.D., Mizuno D.R., Carey S.J. Astron. J. 124, 2990 (2002) Observations of galaxies with the Midcourse Space Experiment.

J/AJ/124/3042/  table1.dat  (14Kb)

Williams R.J., Pogge R.W., Mathur S. Astron. J. 124, 3042 (2002) Narrow-line Seyfert 1 galaxies from the Sloan Digital Sky Survey Early Data Release.

J/AJ/124/3061/  table1.dat table2.dat table3.dat  (18Kb)

Odell A.P., Schombert J., Rakos K. Astron. J. 124, 3061 (2002) The color-magnitude relation in Coma: clues to the age and metallicity of cluster populations.

J/AJ/124/3157/  table3.dat table4.dat  (21Kb)

Valdez-Gutierrez M., Rosado M., Puerari I., Georgiev L., Borissova J., Ambrocio-Cruz P. Astron. J. 124, 3157 (2002) Unveiling the kinematics and dynamics of ionized gas in the nearby irregular galaxy NGC 4449.

J/AJ/124/3222/  table1.dat table2.dat  (748Kb)

Bellazzini M., Ferraro F.R., Origlia L., Pancino E., Monaco L., Oliva E. Astron. J. 124, 3222 (2002) The Draco and Ursa Minor dwarf spheroidal galaxies: a comparative study.

J/AJ/124/3255/  table1.dat  (1.4 Mb)

van der Marel R.P., Gerssen J., Guhathakurta P., Peterson R.C., Gebhardt K. Astron. J. 124, 3255 (2002) Hubble Space Telescope evidence for an intermediate-mass black hole in the globular cluster M15. I. STIS spectroscopy and WFPC2 photometry.

J/AJ/124/3393 See III/240/

J/AJ/124/3465/  table2.dat  (6Kb)

Stevenson S.L., Salzer J.J., Sarajedini V.L., Moran E.C. Astron. J. 124, 3465 (2002) KISS active galactic nuclei in the soft X-ray band: correlation with the ROSAT All-Sky Survey.

J/AJ/125/197/  table1.dat  (15Kb)

Harbeck D., Smith G.H., Grebel E.K. Astron. J., 125, 197-207 (2003) CN abundance variations on the main sequence of 47 Tucanae.

J/AJ/125/277/  table1.dat  (11Kb)

O’dell C.R., Doi T. Astron. J., 125, 277-287 (2003) High proper motion features in the central Orion nebula.

J/AJ/125/359 See III/231/

J/AJ/125/984 See I/284/

J/AJ/125/1261/  table1.dat table2.dat  (86Kb)

Dolphin A.E., Saha A., Skillman E.D., Dohm-Palmer R.C., Tolstoy E., Cole A.A., Gallagher J.S., Hoessel J.G., Mateo M. Astron. J. 125, 1261 (2003) Deep Hubble Space Telescope imaging of Sextans A. II. Cepheids and distance.

J/AJ/125/1298/  table3.dat  (7Kb)

Battinelli P., Demers S., Letarte B. Astron. J. 125, 1298 (2003) Carbon star survey in the Local Group. V. The outer disk of M31.

J/AJ/125/1352/  table1.dat  (85Kb)

Palma C., Majewski S.R., Siegel M.H., Patterson R.J., Ostheimer J.C., Link R. Astron. J. 125, 1352 (2003) Exploring halo substructure with giant stars. IV. The extended structure of the Ursa minor dwarf spheroidal galaxy.

J/AJ/125/1383/  table1.dat  (421Kb)

Twarog B.A., Anthony-Twarog B.J., de Lee N. Astron. J. 125, 1383 (2003) CCD uvbyCaHβphotometry of clusters. III. The most metal-rich open cluster, NGC 6253.

J/AJ/125/1397/  refs.dat table1.dat table2.dat  (30Kb)

Chen L., Hou J.L., Wang J.J. Astron. J., 125, 1397-1406 (2003) On the galactic disk metallicity distribution from open clusters. I. New catalogs and abundance gradient.

J/AJ/125/1537/  table2.dat table3.dat  (126Kb)

Tsujimoto M., Koyama K., Kobayashi N., Goto M., Tsuboi Y., Tokunaga A.T. Astron. J. 125, 1537 (2003) Deep near-infrared observations and identifications of Chandra sources in Orion Molecular Clouds 2 and 3.

J/AJ/125/1598/  table1.dat table3.dat table5.dat  (34Kb)

Lepine S., Rich R.M., Shara M.M. Astron. J. 125, 1598 (2003) Spectroscopy of new high proper motion stars in the northern sky. I. New nearby stars, new high-velocity stars, and an enhanced classification scheme for M dwarfs.

J/AJ/125/1660/  table1.dat table2.dat  (35Kb)

Iovino A., de Carvalho R.R., Gal R.R., Odewahn S.C., Lopes P.A.A., Mahabal A., Djorgovski S.G. Astron. J. 125, 1660 (2003) A new sample of distant compact groups from the Digitized Second Palomar Observatory Sky Survey.

J/AJ/125/1696/  table2.dat table3.dat  (17Kb)

Hancock M., Weistrop D., Eggers D., Nelson C.H. Astron. J. 125, 1696 (2003) Star-forming knots in the UV-bright interacting galaxies NGC 3395 and NGC 3396.

J/AJ/125/1711/  table1.dat table2.dat  (31Kb)

Reichard T.A., Richards G.T., Schneider D.P., Hall P.B., Tolea A., Krolik J.H., Tsvetanov Z., Vanden Berk D.E., York D.G., Knapp G.R., Gunn J.E., Brinkmann J. Astron. J. 125, 1711 (2003) A catalog of Broad Absorption Line Quasars from the Sloan Digital Sky Survey Early Data Release.

J/AJ/125/1817/ table2a.dat table2b.dat table2c.dat table2d.dat table3a.dat table3b.dat table3c.dat table3d.dat (7.2 Mb) Bernardi M., Sheth R.K., Annis J., Burles S., Eisenstein D.J., Finkbeiner D.P., Hogg D.W., Lupton R.H., Schlegel D.J., Subbarao M., Bahcall N.A., Blakeslee J.P., Brinkmann J., Castander F.J., Connolly A.J., Csabai I., Doi M., Fukugita M., Frieman J., Heckman T., Hennessy G.S., Ivezic Z., Knapp G.R., Lamb D.Q., McKay T., Munn J.A., Nichol R., Okamura S., Schneider D.P., Thakar A.R., York D.G. Astron. J. 125, 1817 (2003) Early-type galaxies in the Sloan Digital Sky Survey. I. The sample.

J/AJ/125/1940/  table1.dat table4.dat  (22Kb)

Barba R.H., Rubio M., Roth M.R., Garcia J. Astron. J. 125, 1940 (2003) Active star formation in the N11B nebula in the Large Magellanic Cloud: a sequential star formation scenario confirmed.

J/AJ/125/2064/  table6.dat  (475Kb)

Gal R.R., de Carvalho R.R., Lopes P.A.A., Djorgovski S.G., Brunner R.J., Mahabal A., Odewahn S.C. Astron. J. 125, 2064 (2003)

The northern optical cluster survey. II. An objective cluster catalog for 5800 square degrees.

J/AJ/125/2239/  table1.dat  (14Kb)

O’dwyer I.J., Chu Y.-H., Gruendl R.A., Guerrero M.A., Webbink R.F. Astron. J., 125, 2239-2254 (2003) Hard X-ray emission associated with white dwarfs.

J/AJ/125/2307/  notes.dat table6.dat  (165Kb)

Alonso M.V., Bernardi M., da Costa L.N., Wegner G., Willmer C.N.A., Pellegrini P.S., Maia M.A.G. Astron. J. 125, 2307 (2003) Redshift-distance survey of early-type galaxies: circular-aperture photometry.

J/AJ/125/2393/  table1.dat table2.dat table3.dat  (19Kb)

Miller N.A., Owen F.N., Hill J.M. Astron. J. 125, 2393 (2003) A comprehensive radio and optical study of Abell 2256: activity from an infalling group.

J/AJ/125/2411/  table1.dat table2.dat  (381Kb)

Condon J.J., Cotton W.D., Yin Q.F., Shupe D.L., Storrie-Lombardi L.J., Helou G., Soifer B.T., Werner M.W. Astron. J. 125, 2411 (2003) The SIRTF first-look survey. I. VLA image and source catalog.

J/AJ/125/2427/  table1.dat table2.dat  (18Kb)

Miller N.A., Owen F.N. Astron. J., 125, 2427-2446 (2003) Abell 2255: increased star formation and AGN activity in a cluster-cluster merger.

J/AJ/125/2502/  refs.dat table1.dat table2.dat table4.dat table5.dat table6.dat table7.dat  (42Kb)

Morrison H.L., Norris J., Mateo M., Harding P., Olszewski E.W., Shectman S.A., Dohm-Palmer R.C., Helmi A., Freeman K.C. Astron. J. 125, 2502 (2003) Mapping the galactic halo. VI. Spectroscopic measures of luminosity and metallicity.

J/AJ/125/2521/  table6.dat table7.dat  (8.2 Mb)

Beichman C.A., Cutri R., Jarrett T., Stiening R., Skrutskie M. Astron. J. 125, 2521 (2003) A deep 2MASS survey of the Lockman Hole.

J/AJ/125/2728/  table4.dat  (24Kb)

Assafin M., Zacharias N., Rafferty T.J., Zacharias M.I., Da Silva Neto D.N., Andrei A.H., Vieira Martins R. Astron. J. 125, 2728 (2003) Optical positions of ICRF sources using UCAC reference stars.

J/AJ/125/3037/  table1.dat  (39Kb)

Demers S., Battinelli P., Letarte B. Astron. J., 125, 3037-3045 (2003) Carbon star survey in the Local Group. VI. The dwarf spheroidal galaxy NGC 205.

J/AJ/125/3302/  table1.dat table2.dat  (17Kb)

Gizis J.E., Reid I.N., Knapp G.R., Liebert J., Kirkpatrick J.D., Koerner D.W., Burgasser A.J. Astron. J. 125, 3302 (2003) Hubble Space Telescope observations of binary very low mass stars and brown dwarfs.

J/AJ/125/3311/  table2.dat table3.dat  (66Kb)

Kidger M.R., Martin-Luis F. Astron. J. 125, 3311 (2003) High-precision near-infrared photometry of a large sample of bright stars visible from the northern hemisphere.

J/AJ/126/63/  table1.dat  (16Kb)

Hutchings J.B., Maddox N., Cutri R.M., Nelson B.O. Astron. J. 126, 63 (2003) Host galaxies of 2MASS-selected QSOs to redshift 0.3.

J/AJ/126/119/  table1.dat table2.dat table3.dat table4.dat table5.dat  (41Kb)

Rizza E., Morrison G.E., Owen F.N., Ledlow M.J., Burns J.O., Hill J. Astron. J. 126, 119 (2003) Sensitive radio and optical observations of z 0.2 rich Abell clusters.

J/AJ/126/175/ table2.dat table3.dat table4.dat table5.dat table6.dat table7.dat table8.dat (2.1 Mb) Bonanos A.Z., Stanek K.Z., Sasselov D.D., Mochejska B.J., Macri L.M., Kaluzny J. Astron. J. 126, 175 (2003) DIRECT distances to nearby galaxies using detached eclipsing binaries and Cepheids. IX. Variables in the

field M31Y discovered with image subtraction.

J/AJ/126/237/  table1.dat table2.dat  (124Kb)

Kaluzny J., Rucinski S.M. Astron. J. 126, 237 (2003) Open cluster LW 55 in the Large Magellanic Cloud.

J/AJ/126/472/  table1.dat  (114Kb)

Gould A. Astron. J. 126, 472 (2003) Completeness of USNO-B for high proper motion stars.

J/AJ/126/539/  cdfn-jnl.dat cdfn-sup.dat cdfn.dat cdfs-jnl.dat cdfs-sup.dat cdfs.dat  (329Kb)

Alexander D.M., Bauer F.E., Brandt W.N., Schneider D.P., Hornschemeier A.E., Vignali C., Barger A.J., Broos P.S., Cowie L.L., Garmire G.P., Townsley L.K., Bautz M.W., Chartas G., Sargent W.L.W. Astron. J. 126, 539 (2003) The Chandra Deep Field North Survey. XIII. 2 Ms point-source catalogs.

J/AJ/126/632/  table1a.dat  (67Kb)

Barger A.J., Cowie L.L., Capak P., Alexander D.M., Bauer F.E., Fernandez E., Brandt W.N., Garmire G.P., Hornschemeier A.E. Astron. J. 126, 632 (2003) Optical and infrared properties of the 2Ms Chandra Deep Field North X-ray sources.

J/AJ/126/762/  table1.dat  (475Kb)

Chen A.B.-C., Lu P.K., Mendez R.A., van Altena W.F. Astron. J. 126, 762 (2003) Dark matter: local volume density versus total surface density.

J/AJ/126/1023/  notes.dat table1.dat table2.dat table3.dat  (39Kb)

Harris H.C., Liebert J., Kleinman S.J., Nitta A., Anderson S.F., Knapp G.R., Krzesinski J., Schmidt G., STrauss M.A., Vanden Berk D., Eisenstein D., Hawley S., Margon B., Munn J.A., Silvestri N.M., Smith J.A., Szkody P., Collinge M.J., Dahn C.C., Fan X., Hall P.B., Schneider D.P., Brinkmann J., Burles S., Gunn J.E., Hennessy G.S., Hindsley R., Ivezic Z., Kent S., Lamb D.Q., Lupton R.H., Nichol R.C., Pier J.R., Schlegel D.J., Subbarao M., Uomoto A., Yanny B., York D.G. Astron. J. 126, 1023 (2003) An initial survey of white dwarfs in the Sloan Digital Sky Survey.

J/AJ/126/1183/  table3.dat  (8Kb)

Conselice C.J., Bershady M.A., Dickinson M., Papovich C. Astron. J., 126, 1183-1207 (2003) A direct measurement of major galaxy mergers at z 3.

J/AJ/126/1208/  table1.dat  (21Kb)

Sawicki M., Mallen-Ornelas G. Astron. J. 126, 1208 (2003) Redshifts in the Hubble Deep Field South.

J/AJ/126/1286/  table4.dat  (39Kb)

Lin W., Zhou X., Burstein D., Windhorst R.A., Chen J., Chen W.-P., Jiang Z., Kong X., Ma J., Sun W.-H., Wu H., Xue S., Zhu J. Astron. J. 126, 1286 (2003) Hα+[N II] observations of the H II regions in M81.

J/AJ/126/1362/ table3.dat table4.dat table5.dat table6.dat table7.dat table8.dat table9.dat (201Kb) Brown W.R., Allende Prieto C., Beers T.C., Wilhelm R., Geller M.J., Kenyon S.J., Kurtz M.J. Astron. J. 126, 1362 (2003) The Century Survey Galactic Halo Project. I. Stellar spectral analysis.

J/AJ/126/1423/  id.dat new.dat  (169Kb)

Kraemer K.E., Shipman R.F., Price S.D., Mizuno D.R., Kuchar T., Carey S.J. Astron. J. 126, 1423 (2003) Observations of star-forming regions with the Midcourse Space Experiment.

J/AJ/126/1455/  notes.dat table1.dat table2.dat  (147Kb)

Stark M.A., Wade R.A. Astron. J. 126, 1455 (2003) Single and composite hot subdwarf stars in the light of 2MASS photometry.

J/AJ/126/1526/  table1.dat  (22Kb)

Bouy H., Brandner W., Martin E.L., Delfosse X., Allard F., Basri G.

Astron. J. 126, 1526 (2003) Multiplicity of nearby free-floating ultracool dwarfs: a Hubble Space Telescope WFPC2 search for companions.

J/AJ/126/1607/  table1.dat table2.dat table3.dat  (154Kb)

Sanders D.B., Mazzarella J.M., Kim D.-C., Surace J.A., Soifer B.T. Astron. J. 126, 1607 (2003) The IRAS Revised Bright Galaxy Sample.

J/AJ/126/1665/  table1.dat  (12Kb)

Liu W.M., Meyer M.R., Cotera A.S., Young E.T. Astron. J. 126, 1665 (2003) Hubble Space Telescope NICMOS observations of the embedded cluster in NGC 2024: constraints on the initial mass function and binary fraction .

J/AJ/126/1707/  table1.dat table2.dat  (39Kb)

Garcia-Barreto J.A., Carrillo R., Vera-Villamizar N. Astron. J. 126, 1707 (2003) Companions of bright barred Shapley-Ames galaxies.

J/AJ/126/1720/  table1.dat table2.dat table3.dat table5.dat  (36Kb)

Strateva I.V., Strauss M.A., Hao L., Schlegel D.J., Hall P.B., Gunn J.E., Li L.-X., Ivezic Z., Richards G.T., Zakamska N.L., Voges W., Anderson S.F., Lupton R.H., Schneider D.P., Brinkmann J., Nichol R.C. Astron. J. 126, 1720 (2003) Double-peaked low-ionization emission lines in active galactic nuclei.

J/AJ/126/1750/  table1.dat  (16Kb)

Maia M.A.G., Machado R.S., Willmer C.N.A. Astron. J. 126, 1750 (2003) The Seyfert population in the local universe.

J/AJ/126/1916/  refs.dat table1.dat table2.dat  (34Kb)

Porras A., Christopher M., Allen L., Di Francesco J., Megeath S.T., Myers P.C. Astron. J., 126, 1916-1924 (2003) A catalog of young stellar groups and clusters within 1 kiloparsec of the Sun.

J/AJ/126/1996/  table1.dat table4.dat  (24Kb)

Makarov V.V. Astron. J., 126, 1996-2008 (2003) The 100 brightest X-ray stars within 50 parsecs of the Sun.

J/AJ/126/2048/  notes.dat table1.dat table2.dat table5.dat table7.dat  (91Kb)

Gray R.O., Corbally C.J., Garrison R.F., Mcfadden M.T., Robinson P.E. Astron. J., 126, 2048-2059 (2003) Contributions to the nearby stars (NStars) project: spectroscopy of stars earlier than M0 within 40 parsecs: the Northern Sample. I.

J/AJ/126/2125/  table1.dat  (42Kb)

Zakamska N.L., Strauss M.A., Krolik J.H., Collinge M.J., Hall P.B., Hao L., Heckman T.M., Ivezic Z., Richards G.T., Schlegel D.J., Schneider D.P., Strateva I., Vanden Berk D.E., Anderson S.F., Brinkmann J. Astron. J., 126, 2125-2144 (2003) Candidate type II quasars from the Sloan Digital Sky Survey. I. Selection and optical properties of a sample at 0.3 z 0.83.

J/AJ/126/2152/ table1.dat table3.dat table4.dat table5.dat table6.dat table7.dat table8.dat (856Kb) Rines K., Geller M.J., Kurtz M.J., Diaferio A. Astron. J., 126, 2152-2170 (2003) CAIRNS: the Cluster And Infall Region Nearby Survey. I. Redshifts and mass profiles.

J/AJ/126/2209/  table1.dat table2.dat table3.dat table4.dat table5.dat table6.dat  (243Kb)

Anderson S.F., Voges W., Margon B., Trumper J., Agueros M.A., Boller T., Collinge M.J., Homer L., Stinson G., Strauss M.A., Annis J., Gomez P., Hall P.B., Nichol R.C., Richards G.T., Schneider D.P., Vanden Berk D.E., Fan X., Ivezic Z., Munn J.A., Newberg H.J., Richmond M.W., Weinberg D.H., Yanny B., Bahcall N.A., Brinkmann J., Fukugita M., York D.G. Astron. J., 126, 2209-2229 (2003) A large, uniform sample of X-Ray-Emitting AGNs: selection approach and an initial catalog from the ROSAT All-Sky and Sloan Digital Sky Surveys.

J/AJ/126/2237/  refs.dat table1.dat table3.dat  (54Kb)

Drake C.L., McGregor P.J., Dopita M.A., Van Breugel W.J.M. Astron. J., 126, 2237-2267 (2003) Radio-excess IRAS galaxies: PMN/FSC sample selection.

J/AJ/126/2268/  table1.dat table4.dat table5.dat  (252Kb)

Wegner G., Bernardi M., Willmer C.N.A., Da Costa L.N., Alonso M.V., Pellegrini P.S., Maia M.A.G., Chaves O.L., Rite C. Astron. J., 126, 2268-2280 (2003) Redshift-distance survey of early-type galaxies: spectroscopic data.

J/AJ/126/2421/ notes.dat refs.dat table5.dat table7.dat table8.dat table9.dat table10.dat table11.dat table12.dat table13.dat table14.dat table15.dat table16.dat table17.dat table18.dat table19.dat (110Kb) Cruz K.L., Reid I.N., Liebert J., Kirkpatrick J.D., Lowrance P.J. Astron. J., 126, 2421-2448 (2003) Meeting the cool neighbors. V. A 2MASS-Selected sample of ultracool dwarfs.

J/AJ/126/2449/  table1.dat table2.dat table3.dat  (30Kb)

Reid I.N. Astron. J., 126, 2449-2461 (2003) Meeting the cool neighbors. VI. A search for nearby ultracool dwarfs in the galactic plane.

J/AJ/126/2487/  table1.dat table2.dat  (15Kb)

Burgasser A.J., Mcelwain M.W., Kirkpatrick J.D. Astron. J., 126, 2487-2494 (2003) The 2MASS wide-field T dwarf search. II. Discovery of three T dwarfs in the southern hemisphere.

J/AJ/126/2562/  table3.dat  (34Kb)

Fomalont E.B., Petrov L., Macmillan D.S., Gordon D., Ma C. Astron. J., 126, 2562-2566 (2003) The second VLBA calibrator survey: VCS2.

J/AJ/126/2579/  table1.dat  (10Kb)

Schneider D.P., Fan X., Hall P.B., Jester S., Richards G.T., Stoughton C., Strauss M.A., Subbarao M., Vanden Berk D.E., Anderson S.F., Brandt W.N., Gunn J.E., Gray J., Trump J.R., Voges W., Yanny B., Bahcall N.A., Blanton M.R., Boroski W.N., Brinkmann J., Brunner R., Burles S., Castander F.J., Doi M., Eisenstein D., Frieman J.A., Fukugita M., Heckman T.M., Hennessy G.S., Ivezic Z., Kent S., Knapp G.R., Lamb D.Q., Lee B.C., Loveday J., Lupton R.H., Margon B., Meiksin A., Munn J.A., Newberg H.J., Nichol R.C., Niederste-Ostholt M., Pier J.R., Richmond M.W., Rockosi C.M., Saxe D.H., Schlegel D.J., Szalay A.S., Thakar A.R., Uomoto A., York D.G. Astron. J., 126, 2579-2593 (2003) The Sloan Digital Sky Survey quasar catalog. II. First data release.

J/AJ/126/2644/ cl0024 b.fit cl24 cat.fit cl0024 i.fit cl0024 u.fit cl0024 v.fit cl24ubvi.fit table3.dat (64.7 Mb) Alexov A., Silva D.R., Pierce M.J. Astron. J., 126, 2644-2661 (2003) A wide-field, broadband imaging survey of Butcher-Oemler cluster Cl 0024+1654: the catalog.

J/AJ/126/2740/  table1.dat table2.dat  (106Kb)

Ledlow M.J., Voges W., Owen F.N., Burns J.O. Astron. J., 126, 2740-2751 (2003) The X-ray properties of nearby Abell clusters from the ROSAT All-Sky Survey: the sample and correlations with optical properties.

J/AJ/126/2840/  table3.dat table6.dat  (1.4 Mb)

Lee M.G., Park H.S., Park J.-H., Sohn Y.-J., Oh S.J., Yuk I.-S., Rey S.-C., Lee S.-G., Lee Y.-W., Kim H.-I., Han W., Park W.-K., Lee J.H., Jeon Y.-B., Kim S.C. Astron. J., 126, 2840-2866 (2003) Deep wide-field BVI CCD photometry of the Sextans dwarf spheroidal galaxy.

J/AJ/126/2867/  table1.dat table2.dat table7.dat table8.dat  (35Kb)

Massey P., Olsen K.A.G. Astron. J., 126, 2867-2886 (2003) The evolution of massive stars. I. Red supergiants in the Magellanic Clouds.

J/AJ/126/2922/  table3.dat  (1.1 Mb)

Platais I., Kozhurina-platais V., Mathieu R.D., Girard T.M., Van Altena W.F. Astron. J., 126, 2922-2935 (2003) WIYN open cluster study. XVII. Astrometry and membership to V=21 in NGC 188.

J/AJ/126/2954/  table3.dat table4.dat  (106Kb)

Sandquist E.L., Shetrone M.D. Astron. J., 126, 2954-2962 (2003) S986 in M67: a totally eclipsing binary at the cluster turnoff.

J/AJ/126/2963/  table1.dat  (10Kb)

Muthu C., Anandarao B.G. Astron. J., 126, 2963-2970 (2003) A spatiokinematic study of the planetary nebula NGC 1514.

J/AJ/126/2971/  refs.dat table1.dat table2.dat table5.dat table6.dat  (40Kb)

Vieira S.L.A., Corradi W.J.B., Alencar S.H.P., Mendes L.T.S., Torres C.A.O., Quast G.R., Guimaraes M.M., Da Silva L. Astron. J., 126, 2971-2987 (2003) Investigation of 131 Herbig Ae/Be candidate stars.

J/AJ/126/3007/  notes.dat table1.dat table2.dat table3.dat table4.dat  (88Kb)

Reid I.N., Cruz K.L., Allen P., Mungall F., Kilkenny D., Liebert J., Hawley S.L., Fraser O.J., Covey K.R., Lowrance P. Astron. J., 126, 3007-3016 (2003) Meeting the cool neighbors. VII. Spectroscopy of faint red NLTT dwarfs.

J/AJ/126/3017/  table2.dat table4.dat table5.dat  (123Kb)

Thorstensen J.R. Astron. J., 126, 3017-3029 (2003) Parallaxes and distance estimates for 14 cataclysmic variable stars.

J/AJ/127/24/  table4.dat table5.dat table6.dat table7.dat  (50Kb)

Jordan A., Cote P., West M.J., Marzke R.O., Minniti D., Rejkuba M. Astron. J., 127, 24-47 (2004) Hubble Space Telescope observations of cD galaxies and their globular cluster systems.

J/AJ/127/105/  table1.dat  (8Kb)

Boker T., Sarzi M., Mclaughlin D.E., Van Der Marel R.P., Rix H.-W., Ho L.C., Shields J.C. Astron. J., 127, 105-118 (2004) A Hubble Space Telescope census of nuclear star clusters in late-type spiral galaxies. II. Cluster sizes and structural parameter correlations.

J/AJ/127/131/  table1.dat  (28Kb)

Stanford S.A., Dickinson M., Postman M., Ferguson H.C., Lucas R.A., Conselice C.J., Budavari T., Somerville R. Astron. J., 127, 131-155 (2004) The evolution of early-type field galaxies selected from a NICMOS map of the Hubble Deep Field North.

J/AJ/127/156/  table1.dat table2.dat table3.dat  (36Kb)

Grupe D., Wills B.J., Leighly K.M., Meusinger H. Astron. J., 127, 156-179 (2004) A complete sample of soft X-Ray-Selected AGNs. I. The data.

J/AJ/127/180/ flagr.dat flagz.dat fluxr.dat fluxz.dat magr.dat magz.dat shaper.dat shapez.dat (54.6 Mb) Capak P., Cowie L.L., Hu E.M., Barger A.J., Dickinson M., Fernandez E., Giavalisco M., Komiyama Y., Kretchmer C., Mcnally C., Miyazaki S., Okamura S., Stern D. Astron. J., 127, 180-198 (2004) A deep wide-field, optical, and near-infrared catalog of a large area around the Hubble Deep Field North.

J/AJ/127/213/  table1.dat  (39Kb)

Wang J.X., Malhotra S., Rhoads J.E., Brown M.J.I., Dey A., Heckman T.M., Jannuzi B.T., Norman C.A., Tiede G.P., Tozzi P. Astron. J., 127, 213-229 (2004) The 172 ks Chandra exposure of the LALA Bootes field: X-ray source catalog.

J/AJ/127/318/  refs.dat table2.dat table3.dat table4.dat table5.dat  (49Kb)

Pritzl B.J., Armandroff T.E., Jacoby G.H., Da Costa G.S. Astron. J., 127, 318-333 (2004) The dwarf spheroidal companions to M31: variable stars in Andromeda II.

J/AJ/127/334/  table1.dat table2.dat table3.dat table4.dat table5.dat table6.dat  (218Kb)

Alcock C., Alves D.R., Axelrod T.S., Becker A.C., Bennett D.P., Clement C.M., Cook K.H., Drake A.J., Freeman K.C., Geha M., Griest K., Lehner M.J., Marshall S.L., Minniti D., Muzzin A., Nelson C.A., Peterson B.A., Popowski P., Quinn P.J., Rodgers A.W., Rowe J.F., Sutherland W., Vandehei T., Welch D.L. Astron. J., 127, 334-354 (2004) The MACHO project Large Magellanic Cloud variable-star inventory. XIII. Fourier parameters for the firstovertone RR Lyrae variables and the LMC distance.

J/AJ/127/373/  table1.dat  (9Kb)

Tautvaisiene G., Wallerstein G., Geisler D., Gonzalez G., Charbonnel C. Astron. J., 127, 373-379 (2004) Elemental abundances of three red giants in Terzan 7, a globular cluster associated with the Sagittarius galaxy.

J/AJ/127/477/  table1.dat table1.pos  (14Kb)

Chu Y.-H., Guerrero M.A., Gruendl R.A., Webbink R.F. Astron. J., 127, 477-480 (2004) Hard X-ray emission associated with white dwarfs. II.

J/AJ/127/501/  table1.dat table2.dat  (17Kb)

Lewis B.M., Kopon D.A., Terzian Y. Astron. J., 127, 501-512 (2004) 2MASS counterparts for OH/IR stars. I. The Arecibo sample.

J/AJ/127/539/  table1.dat table2.dat table3.dat table4.dat table5.dat  (135Kb)

Mason B.D., Hartkopf W.I., Wycoff G.L., Pascu D., Urban S.E., Hall D.M., Hennessy G.S., Rafferty T.J., Flagg L., Kang D., Ries P., Holdenried E. Astron. J., 127, 539-548 (2004) Speckle interferometry at the US naval observatory. IX.

J/AJ/127/646/  table5.dat  (75Kb)

Francis P.J., Nelson B.O., Cutri R.M. Astron. J., 127, 646-655 (2004) An unbiased census of active galactic nuclei in the Two Micron All Sky Survey.

J/AJ/127/660/  table1.dat table2.dat  (301Kb)

Saviane I., Hibbard J.E., Rich R.M. Astron. J., 127, 660-678 (2004) The stellar content of the southern tail of NGC 4038/4039 and a revised distance

J/AJ/127/704/  table1.dat table2.dat table3.dat table4.dat table5.dat table6.dat  (67Kb)

Kniazev A.Y., Grebel E.K., Pustilnik S.A., Pramskij A.G., Kniazeva T.F., Prada F., Harbeck D. Astron. J., 127, 704-727 (2004) Low surface brightness galaxies in the Sloan Digital Sky Survey. I. Search method and test sample.

J/AJ/127/840/  table2.dat  (18Kb)

Pont F., Zinn R., Gallart C., Hardy E., Winnick R. Astron. J., 127, 840-860 (2004) The chemical enrichment history of the Fornax dwarf spheroidal galaxy from the infrared calcium triplet.

J/AJ/127/861/  table1.dat table2.dat  (20Kb)

Bonanos A.Z., Stanek K.Z., Szentgyorgyi A.H., Sasselov D.D., Bakos G.A. Astron. J., 127, 861-867 (2004) The RR Lyrae distance to the Draco dwarf spheroidal galaxy.

J/AJ/127/899/  table3.dat  (229Kb)

Sirko E., Goodman J., Knapp G.R., Brinkmann J., Ivezic Z., Knerr E.J., Schlegel D., Schneider D.P., York D.G. Astron. J., 127, 899-913 (2004) Blue horizontal-branch stars in the Sloan Digital Sky Survey. I. Sample selection and structure in the galactic halo.

J/AJ/127/958/  table2.dat table3.dat  (4.2 Mb)

Rey S.-C., Lee Y.-W., Ree C.H., Joo J.-M., Sohn Y.-J., Walker A.R. Astron. J., 127, 958-979 (2004) CCD photometry of the globular cluster ωCentauri. II. Stellar populations and age-metallicity relation.

J/AJ/127/980/  position.dat table2.dat  (252Kb)

Hughes J., Wallerstein G., Van Leeuwen F., Hilker M. Astron. J., 127, 980-990 (2004) The giant branches of ωCentauri: multiwavelength observations of evolved stars.

J/AJ/127/1000/  table2.dat  (61Kb)

Anthony-Twarog B.J., Twarog B.A. Astron. J., 127, 1000-1013 (2004) uvbyCaHβCCD photometry of clusters. IV. Solving the riddle of NGC 3680.

J/AJ/127/1014/  table2.dat table3.dat table4.dat table5.dat  (26Kb)

Sung H., Bessell M.S. Astron. J., 127, 1014-1028 (2004) The initial mass function and stellar content of NGC 3603.

J/AJ/127/1117/  table2.dat  (40Kb)

Reipurth B., Pettersson B., Armond T., Bally J., Vaz L.P.R. Astron. J., 127, 1117-1130 (2004) Hαemission-line stars in molecular clouds. I. The NGC 2264 region.

J/AJ/127/1131/  table1.dat  (22Kb)

Wilking B.A., Meyer M.R., Greene T.P., Mikhail A., Carlson G. Astron. J., 127, 1131-1146 (2004) Low-mass stars and substellar objects in the NGC 1333 molecular cloud.

J/AJ/127/1158/  table2.dat  (48Kb)

Vivas A.K., Zinn R., Abad C., Andrews P., Bailyn C., Baltay C., Bongiovanni A., Briceno C., Bruzual G., Coppi P., Della Prugna F., Ellman N., Ferrin I., Gebhard M., Girard T., Hernandez J., Herrera D., Honeycutt R., Magris G., Mufson S., Musser J., Naranjo O., Rabinowitz D., Rengstorf A., Rosenzweig P., Sanchez Ge., Sanchez Gu., Schaefer B., Schenner H., Snyder J.A., Sofia S., Stock J., Van Altena W., Vicente B., Vieira K. Astron. J., 127, 1158-1175 (2004) The QUEST RR Lyrae survey. I. The first catalog.

J/AJ/127/1181/  table3.dat  (13Kb)

Docobo J.A., Andrade M., Ling J.F., Prieto C., Tamazian V.S., Balega Y.Y., Blanco J., Maximov A.F., Lahulla J.F., Alvarez C. Astron. J., 127, 1181-1186 (2004) Binary star speckle interferometry: measurements and orbits.

J/AJ/127/1227/  table3.dat  (34Kb)

Clem J.L., Vandenberg D.A., Grundahl F., Bell R.A. Astron. J., 127, 1227-1256 (2004) Empirically constrained color-temperature relations. II. uvby.

J/AJ/127/1360/  table2.dat  (8Kb)

Weistrop D., Eggers D., Hancock M., Nelson C.H., Bachilla R., Kaiser M.E. Astron. J., 127, 1360-1370 (2004) Looking closely at medusa: star-forming knots at the center of NGC 4194.

J/AJ/127/1513/  table1.dat table2.dat table3.dat table6.dat  (25Kb)

Schiavon R.P., Caldwell N., Rose J.A. Astron. J., 127, 1513-1530 (2004) The integrated spectrum of M67 and the spectroscopic age of M32.

J/AJ/127/1531/  table2.dat  (707Kb)

Harris J., Zaritsky D. Astron. J., 127, 1531-1544 (2004) The star formation history of the Small Magellanic Cloud.

J/AJ/127/1545/  table3.dat table4.dat  (21Kb)

Cohen J.G. Astron. J., 127, 1545-1554 (2004) Palomar 12 as a part of the Sagittarius stream: the evidence from abundance ratios.

J/AJ/127/1567/  table4.dat  (58Kb)

Parker J.E., Humphreys R.M., Beers T.C. Astron. J., 127, 1567-1578 (2004) The asymmetric thick disk: a star-count and kinematic analysis. II. The kinematics.

J/AJ/127/1579/  table3.dat table4.dat table5.dat  (19Kb)

Briley M.M., Cohen J.G., Stetson P.B. Astron. J., 127, 1579-1587 (2004) The chemical inhomogeneity of faint M13 stars: carbon and nitrogen abundances.

J/AJ/127/1588/  table1.dat  (13Kb)

Briley M.M., Harbeck D., Smith G.H., Grebel E.K. Astron. J., 127, 1588-1593 (2004) On the carbon and nitrogen abundances of 47 Tucanae’s main-sequence stars.

J/AJ/127/1602/  table2.dat table3.dat  (21Kb)

Cohen R.E., Herbst W., Williams E.C. Astron. J., 127, 1602-1621 (2004) A multiyear photometric study of IC 348.

J/AJ/127/1632/  table2.dat table3.dat table4.dat table5.dat  (158Kb)

Oey M.S., King N.L., Parker J.W.

Astron. J., 127, 1632-1643 (2004) Massive field stars and the stellar clustering law.

J/AJ/127/1727/  stars.dat table1.dat  (62Kb)

Horch E.P., Meyer R.D., Van Altena W.F. Astron. J., 127, 1727-1735 (2004) Speckle observations of binary stars with the WIYN telescope. IV. Differential photometry.

J/AJ/127/1811/  table1.dat table2.dat  (101Kb)

Lee B.C., Allam S.S., Tucker D.L., Annis J., Johnston D.E., Scranton R., Acebo Y., Bahcall N.A., Bartelmann M., Bohringer H., Ellman N., Grebel E.K., Infante L., Loveday J., Mckay T.A., Prada F., Schneider D.P., Stoughton C., Szalay A.S., Vogeley M.S., Voges W., Yanny B. Astron. J., 127, 1811-1859 (2004) A catalog of compact groups of galaxies in the SDSS commissioning data.

J/AJ/127/1883/  table2.dat  (199Kb)

Allam S.S., Tucker D.L., Smith J.A., Lee B.C., Annis J., Lin H., Karachentsev I.D., Laubscher B.E. Astron. J., 127, 1883-1899 (2004) Merging galaxies in the Sloan Digital Sky Survey Early Data Release.

J/AJ/127/1943/  table2.dat table4.dat  (100Kb)

Gronwall C., Salzer J.J., Sarajedini V.L., Jangren A., Chomiuk L., Moody J.W., Frattare L.M., Boroson T.A. Astron. J., 127, 1943-1958 (2004) The KPNO International Spectroscopic Survey. IV. Hα-selected survey list 2.

J/AJ/127/1959/  table3.dat  (23Kb)

Van Duyne J., Beckerman E., Salzer J.J., Gronwall C., Thuan T.X., Condon J.J., Frattare L.M. Astron. J., 127, 1959-1976 (2004) Radio continuum emission at 1.4 GHz from KISS emission-line galaxies.

J/AJ/127/1982/  table3.dat  (10Kb)

Buta R.J., Byrd G.G., Freeman T. Astron. J., 127, 1982-2001 (2004) A Hubble space telescope study of star formation in the inner resonance ring of NGC 3081.

J/AJ/127/2002/  table1.dat  (16Kb)

Kewley L.J., Geller M.J., Jansen R.A. Astron. J., 127, 2002-2030 (2004) [O II] as a star formation rate indicator.

J/AJ/127/2031/  table1.dat table3.dat table4.dat  (135Kb)

Karachentsev I.D., Karachentseva V.E., Huchtmeier W.K., Makarov D.I. Astron. J., 127, 2031-2068 (2004) A catalog of neighboring galaxies.

J/AJ/127/2085/  table2.dat  (30Kb)

Anderson K.S.J., Baggett S.M., Baggett W.E. Astron. J., 127, 2085-2093 (2004) Inner-truncated disks in galaxies.

J/AJ/127/2114/  table3.dat table4.dat table5.dat  (65Kb)

Dirsch B., Richtler T., Geisler D., Gebhardt K., Hilker M., Alonso M.V., Forte J.C., Grebel E.K., Infante L., Larsen S., Minniti D., Rejkuba M. Astron. J., 127, 2114-2132 (2004) The globular cluster system of NGC 1399. III. VLT spectroscopy and database.

J/AJ/127/2210/  table4.dat table5.dat table6.dat  (565Kb)

Rider C.J., Tucker D.L., Smith J.A., Stoughton C., Allam S.S., Neilsen E.H.JR Astron. J., 127, 2210-2227 (2004) A survey of open clusters in the u’g’r’i’z’ filter system. I. Results for NGC 2548 (M48).

J/AJ/127/2228/  table3.dat table4.dat  (25Kb)

Makidon R.B., Rebull L.M., Strom S.E., Adams M.T., Patten B.M. Astron. J., 127, 2228-2245 (2004) Periodic variability of pre-main-sequence stars in the NGC 2264 OB association.

J/AJ/127/2455/  table4.dat table5.dat  (69Kb)

Abraham R.G., Glazebrook K., Mccarthy P.J., Crampton D., Murowinski R., Jorgensen I., Roth K., Hook I.M., Savaglio S., Chen H.-W., Marzke R.O., Carlberg R.G. Astron. J., 127, 2455-2483 (2004) The Gemini Deep Deep Survey. I. Introduction to the survey, catalogs, and composite spectra.

J/AJ/127/2484/  tables.dat  (45Kb)

Holden B.P., Stanford S.A., Eisenhardt P., Dickinson M. Astron. J., 127, 2484-2510 (2004) Evolution in the color-magnitude relation of early-type galaxies in clusters of galaxies at z= 1.

J/AJ/127/2565/  table2.dat  (16Kb)

de Vries W.H., Becker R.H., White R.L., Helfand D.J. Astron. J., 127, 2565-2578 (2004) Optical properties of faint FIRST variable radio sources.

J/AJ/127/2579/  table2.dat  (24Kb)

Zhang H., Xue S., Burstein D., Zhou X., Jiang Z., Wu H., Ma J., Chen J., Zou Z. Astron. J., 127, 2579-2597 (2004) Multicolor photometric observations of candidate optical counterparts to ROSAT faint X-ray sources in a 1 square degree field of the BATC Survey.

J/AJ/127/2659/  table1.dat table4.dat table5.dat  (82Kb)

Ramirez S.V., Rebull L., Stauffer J., Hearty T., Hillenbrand L., Jones B., Makidon R., Pravdo S., Strom S., Werner M. Astron. J., 127, 2659-2673 (2004) CHANDRA X-ray observations of young clusters. I. NGC 2264 data.

J/AJ/127/2674/  table1.dat table3.dat table4.dat  (97Kb)

Olsen K.A.G., Miller B.W., Suntzeff N.B., Schommer R.A., Bright J. Astron. J., 127, 2674-2693 (2004) The globular cluster systems of the Sculptor group.

J/AJ/127/2723/  table7.dat  (17Kb)

Kaufer A., Venn K.A., Tolstoy E., Pinte C., Kudritzki R.-P. Astron. J., 127, 2723-2737 (2004) First stellar abundances in the dwarf irregular galaxy Sextans A.

J/AJ/127/2738/  table1.dat table2.dat table3.dat  (377Kb)

Brown T.M., Ferguson H.C., Smith E., Kimble R.A., Sweigart A.V., Renzini A., Rich R.M. Astron. J., 127, 2738-2752 (2004) RR Lyrae stars in the Andromeda halo from deep imaging with the Advanced Camera for Surveys.

J/AJ/127/2771/  table3.dat table4.dat table5.dat table6.dat table7.dat  (144Kb)

Richer H.B., Fahlman G.G., Brewer J., Davis S., Kalirai J., Stetson P.B., Hansen B.M.S., Rich R.M., Ibata R.A., Gibson B.K., Shara M. Astron. J., 127, 2771-2792 (2004) Hubble space telescope observations of the main sequence of M4.

J/AJ/127/2838/  table1.dat  (33Kb)

Downes R.A., Margon B., Anderson S.F., Harris H.C., Knapp G.R., Schroeder J., Schneider D.P., York D.G., Pier J.R., Brinkmann J. Astron. J., 127, 2838-2849 (2004) Faint high-latitude carbon stars discovered by the Sloan Digital Sky Survey: an initial catalog.

J/AJ/127/2856/  refs.dat table2.dat table3.dat  (21Kb)

Burgasser A.J., Mcelwain M.W., Kirkpatrick J.D., Cruz K.L., Tinney C.G., Reid I.N. Astron. J., 127, 2856-2870 (2004) The 2MASS wide-field T dwarf search. III. Seven new T dwarfs and other cool dwarf discoveries.

J/AJ/127/2915/  notes.dat table1.dat table2.dat table3.dat table5.dat  (62Kb)

Parsons S.B. Astron. J., 127, 2915-2930 (2004) New and confirmed triple systems with luminous cool primaries and hot companions

J/AJ/127/3023/  table2.dat  (10Kb)

Sheppard S.S., Jewitt D. Astron. J., 127, 3023-3033 (2004) Extreme Kuiper belt object 2001 QG_{298} and the fraction of contact binaries.

J/AJ/127/3043 See I/289/

J/AJ/127/3060 See I/288/

J/AJ/127/3075/  table2.dat table3.dat  (73Kb)

La Franca F., Gruppioni C., Matute I., Pozzi F., Lari C., Mignoli M., Zamorani G., Alexander D.M., Cocchia F., Danese L., Franceschini A., Heraudeau P., Kotilainen J.K., Linden-Vornle M.J.D., Oliver S., Rowan-Robinson M., Serjeant S., Spinoglio L., Verma A.

Astron. J., 127, 3075-3088 (2004) The nature of the mid-infrared population from optical identifications of the ELAIS-S1 sample.

J/AJ/127/3121/  catalog.dat refs.dat table1.dat  (770Kb)

Wirth G.D., Willmer C.N.A., Amico P., Chaffee F.H., Goodrich R.W., Kwok S., Lyke J.E., Mader J.A., Tran H.D., Barger A.J., Cowie L.L., Capak P., Coil A.L., Cooper M.C., Conrad A., Davis M., Faber S.M., Hu E.M., Koo D.C., Le Mignant D., Newman J.A., Songaila A. Astron. J., 127, 3121-3136 (2004) The Team Keck Treasury Redshift Survey of the GOODS-North Field.

J/AJ/127/3137/  refs.dat table1.dat table2.dat table3.dat  (927Kb)

Cowie L.L., Barger A.J., Hu E.M., Capak P., Songaila A. Astron. J., 127, 3137-3145 (2004) A large sample of spectroscopic redshifts in the ACS-GOODS region of the Hubble Deep Field North.

J/AJ/127/3213/  table1.dat table3.dat tablea1.dat  (30Kb)

Marcum P.M., Aars C.E., Fanelli M.N. Astron. J., 127, 3213-3234 (2004) Early-type galaxies in extremely isolated environments: typical ellipticals.

J/AJ/127/3235/  table1.dat  (38Kb)

Surace J.A., Sanders D.B., Mazzarella J.M. Astron. J., 127, 3235-3272 (2004) An IRAS high resolution image restoration (HIRES) atlas of all interacting galaxies in the IRAS Revised Bright Galaxy Sample.

J/AJ/127/3273/  table2.dat table3.dat table4.dat  (101Kb)

Vogt N.P., Haynes M.P., Herter T., Giovanelli R. Astron. J., 127, 3273-3299 (2004) M/L, Hαrotation curves, and H I measurements for 329 nearby cluster and field spirals. I. Data.

J/AJ/127/3399/  table2.dat  (13Kb)

Herrnstein R.M., Zhao J.-H., Bower G.C., Goss W.M. Astron. J., 127, 3399-3410 (2004) The variability of Sagittarius A* at centimeter wavelengths.

J/AJ/127/3456/  table1.dat table2.dat  (25Kb)

Doi T., O’Dell C.R., Hartigan P. Astron. J., 127, 3456-3478 (2004) Internal velocities in the Orion nebula: large radial velocity features.

J/AJ/127/3553/  table1.dat table3.dat table9.dat  (30Kb)

Knapp G.R., Leggett S.K., Fan X., Marley M.S., Geballe T.R., Golimowski D.A., Finkbeiner D., Gunn J.E., Hennawi J., Ivezic Z., Lupton R.H., Schlegel D.J., Strauss M.A., Tsvetanov Z.I., Chiu K., Hoversten E.A., Glazebrook K., Zheng W., Hendrickson M., Williams C.C., Uomoto A., Vrba F.J., Henden A.A., Luginbuhl C.B., Guetter H.H., Munn J.A., Canzian B., Schneider D.P., Brinkmann J. Astron. J., 127, 3553-3578 (2004) Near-infrared photometry and spectroscopy of L and T dwarfs: the effects of temperature, clouds, and gravity.

J/AJ/127/3579/  stars.dat table1.dat  (45Kb)

Paulson D.B., Cochran W.D., Hatzes A.P. Astron. J., 127, 3579-3586 (2004) Searching for planets in the Hyades. V. Limits on planet detection in the presence of stellar activity.

J/AJ/127/3587/  table1.dat table2.dat table3.dat  (68Kb)

Fey A.L., Ma C., Arias E.F., Charlot P., Feissel-Vernier M., Gontier A.-M., Jacobs C.S., Li J., Macmillan D.S. Astron. J., 127, 3587-3608 (2004) The second extension of the International Celestial Reference Frame: ICRF-EXT.1.

J/AJ/127/3609/  objects.dat table2.dat table3.dat table4.dat  (24Kb)

Ojha R., Fey A.L., Johnston K.J., Jauncey D.L., Reynolds J.E., Tzioumis A.K., Quick J.F.H., Nicolson G.D., Ellingsen S.P., Dodson R.G., Mcculloch P.M. Astron. J., 127, 3609-3621 (2004) VLBI observations of southern hemisphere ICRF sources. I.

J/AJ/128/1/  elais/* fls/* table2.dat table5.dat  (26.9 Mb)

Fadda D., Jannuzi B.T., Ford A., Storrie-Lombardi L.J. Astron. J., 128, 1-15 (2004) The Spitzer Space Telescope First-Look Survey: KPNO Mosaic-1 R-band images and source catalogs.

J/AJ/128/16/  table2.dat  (133Kb)

Koribalski B.S., Staveley-smith L., Kilborn V.A., Ryder S.D., Kraan-Korteweg R.C., Ryan-weber E.V., Ekers R.D., Jerjen H., Henning P.A., Putman M.E., Zwaan M.A., De Blok W.J.G., Calabretta M.R., Disney M.J., Minchin R.F., Bhathal R., Boyce P.J., Drinkwater M.J., Freeman K.C., Gibson B.K., Green A.J., Haynes R.F., Juraszek S., Kesteven M.J., Knezek P.M., Mader S., Marquarding M., Meyer M., Mould J.R., Oosterloo T., O’Brien J., Price R.M., Sadler E.M., Schroeder A., Stewart I.M., Stootman F., Waugh M., Warren B.E., Webster R.L., Wright A.E. Astron. J., 128, 16-46 (2004) The 1000 brightest HIPASS galaxies: H I properties.

J/AJ/128/62/  table1.dat  (14Kb)

Gimeno G.N., Diaz R.J., Carranza G.J. Astron. J., 128, 62-67 (2004) Catalog of double nucleus disk galaxies.

J/AJ/128/163/  notes.dat table2.dat table3.dat table5.dat table7.dat  (44Kb)

Lotz J.M., Primack J., Madau P. Astron. J., 128, 163-182 (2004) A new nonparametric approach to galaxy morphological classification.

J/AJ/128/224/  table4.dat  (1.0 Mb)

Tiede G.P., Sarajedini A., Barker M.K. Astron. J., 128, 224-236 (2004) The stellar populations in the outer regions of M33. I. Metallicity distribution function.

J/AJ/128/245/  table3.dat  (31Kb)

Majewski S.R., Kunkel W.E., Law D.R., Patterson R.J., Polak A.A., Rocha-Pinto H.J., Crane J.D., Frinchaboy P.M., Hummels C.B., Johnston K.V., Rhee J., Skrutskie M.F., Weinberg M. Astron. J., 128, 245-259 (2004) A Two Micron All Sky Survey view of the Sagittarius dwarf galaxy. II. Swope telescope spectroscopy of M giant stars in the dynamically cold Sagittarius tidal stream.

J/AJ/128/343/  stars.dat table2.dat tablea1.dat tablea2.dat tablea3.dat  (674Kb)

Luck R.E., Andrievsky S.M. Astron. J., 128, 343-356 (2004) Phase-dependent variation of the fundamental parameters of cepheids. I. Periods from 6 to 10 days.

J/AJ/128/426/  table2.dat  (9Kb)

West A.A., Hawley S.L., Walkowicz L.M., Covey K.R., Silvestri N.M., Raymond S.N., Harris H.C., Munn J.A., Mcgehee P.M., Ivezic Z., Brinkmann J. Astron. J., 128, 426-436 (2004) Spectroscopic properties of cool stars in the Sloan Digital Sky Survey: an analysis of magnetic activity and a search for subdwarfs.

J/AJ/128/463/ notes.dat refs.dat table1.dat table2.dat table3.dat table4.dat table5.dat table6.dat table7.dat table8.dat (184Kb) Reid I.N., Cruz K.L., Allen P., Mungall F., Kilkenny D., Liebert J., Hawley S.L., Fraser O.J., Covey K.R., Lowrance P., Kirkpatrick J.D., Burgasser A.J. Astron. J., 128, 463-483 (2004) Meeting the cool neighbors. VIII. A preliminary 20 parsec census from the NLTT catalogue.

J/AJ/128/544/  table5.dat table6.dat  (12Kb)

McIntosh D.H., Impey C.D., Petry C.E. Astron. J., 128, 544-560 (2004) Quasars as absorption probes of the J0053+1234 region.

J/AJ/128/712/  table3.dat  (8.2 Mb)

Harris G.L.H., Geisler D., Harris W.E., Schmidt B.P., Hesser J.E., Reid M., Milne M., Hulme S.C., Kidd T.T. Astron. J., 128, 712-722 (2004) Wide-field Washington photometry of the NGC 5128 globular cluster system. I. The database.

J/AJ/128/723/  table2.dat table4.dat  (35Kb)

Harris G.L.H., Harris W.E., Geisler D. Astron. J., 128, 723-735 (2004) Wide-field Washington photometry of the NGC 5128 globular cluster system. II. Large-scale properties of the system.

J/AJ/128/787/  table1.dat table2.dat table3.dat table4.dat  (115Kb)

Ramirez S.V., Rebull L., Stauffer J., Strom S., Hillenbrand L., Hearty T., Kopan E.L., Pravdo S., Makidon R., Jones B. Astron. J., 128, 787-804 (2004)

Chandra X-ray observations of young clusters. II. Orion flanking fields data.

J/AJ/128/805/  table2.dat table3.dat table6.dat table7.dat  (21Kb)

Sicilia-Aguilar A., Hartmann L.W., Briceno C., Muzerolle J., Calvet N. Astron. J., 128, 805-821 (2004) Low-mass stars and accretion at the ages of planet formation in the Cepheus OB2 region.

J/AJ/128/846/  table2.dat table3.dat  (17Kb)

Kang Y.W., Lee H.-W., Hong K.S., Kim C.-H., Guinan E.F. Astron. J., 128, 846-857 (2004) The chromospherically active contact binary CE Leonis.

J/AJ/128/1002/  table1.dat table2.dat table3.dat table4.dat  (26Kb)

Zakamska N.L., Strauss M.A., Heckman T.M., Ivezic Z., Krolik J.H. Astron. J., 128, 1002-1016 (2004) Candidate type II quasars from the Sloan Digital Sky Survey. II. From radio to x-rays.

J/AJ/128/1017/  table1.dat  (689Kb)

Lopes P.A.A., De Carvalho R.R., Gal R.R., Djorgovski S.G., Odewahn S.C., Mahabal A.A., Brunner R.J. Astron. J., 128, 1017-1045 (2004) + Erratum 130, 2434 (2005) The northern sky optical cluster survey. IV. An intermediate-redshift galaxy cluster catalog and the comparison of two detection algorithms.

J/AJ/128/1078/  table4.dat  (22Kb)

Rines K., Geller M.J., Diaferio A., Kurtz M.J., Jarrett T.H. Astron. J., 128, 1078-1111 (2004) CAIRNS: the cluster and infall region nearby survey. II. Environmental dependence of infrared mass-to-light ratios.

J/AJ/128/1167/  table1.dat table2.dat  (119Kb)

Gieren W., Pietrzynski G., Walker A., Bresolin F., Minniti D., Kudritzki R.-P., Udalski A., Soszynski I., Fouque P., Storm J., Bono G. Astron. J., 128, 1167-1176 (2004) The Araucaria project: an improved distance to the Sculptor spiral galaxy NGC 300 from its Cepheid variables.

J/AJ/128/1177/  refs.dat table2.dat  (122Kb)

Venn K.A., Irwin M., Shetrone M.D., Tout C.A., Hill V., Tolstoy E. Astron. J., 128, 1177-1195 (2004) Stellar chemical signatures and hierarchical galaxy formation.

J/AJ/128/1233/  table1.dat table4.dat table8.dat  (53Kb)

Herbig G.H., Andrews S.M., Dahm S.E. Astron. J., 128, 1233-1253 (2004) LkHα101 and the young cluster in NGC 1579.

J/AJ/128/1254/  table2.dat  (56Kb)

Lada C.J., Muench A.A., Lada E.A., Alves J.F. Astron. J., 128, 1254-1264 (2004) Deep 3.8 micron observations of the Trapezium cluster.

J/AJ/128/1319/  table1.dat  (81Kb)

Williamon R.M., Sowell J.R., Van Hamme W. Astron. J., 128, 1319-1323 (2004) Orbital solutions and absolute elements of the eclipsing binary AY Camelopardalis.

J/AJ/128/1324/  table3.dat  (118Kb)

Lacy C.H.S., Vaz L.P.R., Claret A., Sabby J.A. Astron. J., 128, 1324-1330 (2004) Absolute properties of the main-sequence eclipsing binary star V885 Cygni.

J/AJ/128/1340/  table5.dat  (142Kb)

Lacy C.H.S., Claret A., Sabby J.A. Astron. J., 128, 1340-1347 (2004) Absolute properties of the eclipsing binary star V459 Cassiopeiae.

J/AJ/128/1430 See I/283A/

J/AJ/128/1483/  table1.dat  (68Kb)

Steffen A.T., Barger A.J., Capak P., Cowie L.L., Mushotzky R.F., Yang Y. Astron. J., 128, 1483-1500 (2004) An optical catalog of the Chandra Large Area Synoptic X-ray Survey sources.

J/AJ/128/1501/  table2.dat table3.dat  (129Kb)

Yang Y., Mushotzky R.F., Steffen A.T., Barger A.J., Cowie L.L. Astron. J., 128, 1501-1523 (2004) The Chandra Large Area Synoptic X-ray Survey (CLASXS) of the Lockman Hole-Northwest: the X-ray catalog.

J/AJ/128/1558/  table1.dat table5.dat  (481Kb)

Smith R.J., Hudson M.J., Nelan J.E., Moore S.A.W., Quinney S.J., Wegner G.A., Lucey J.R., Davies R.L., Malecki J.J., Schade D., Suntzeff N.B. Astron. J., 128, 1558-1569 (2004) NOAO Fundamental Plane Survey. I. Survey design, redshifts, and velocity dispersion data.

J/AJ/128/1570/  table1.dat table2.dat  (13Kb)

Ojha R., Fey A.L., Lovell J.E.J., Jauncey D.L., Johnston K.J. Astron. J., 128, 1570-1587 (2004) VLBA snapshot imaging survey of scintillating sources.

J/AJ/128/1646/  table2.dat table3.dat table4.dat table5.dat table6.dat  (49Kb)

Nord M.E., Lazio T.J.W., Kassim N.E., Hyman S.D., Larosa T.N., Brogan C.L., Duric N. Astron. J., 128, 1646-1670 (2004) High-resolution, wide-field imaging of the Galactic Center region at 330 MHz.

J/AJ/128/1684/  table1.dat table3.dat  (393Kb)

Sung H., Bessell M.S., Chun M.-Y. Astron. J., 128, 1684-1693 (2004) The initial mass function and young brown dwarf candidates in NGC 2264. I. The initial mass function around S Monocerotis.

J/AJ/128/1748/  table1.dat table2.dat  (67Kb)

Schmidt E.G., Johnston D., Langan S., Lee K.M. Astron. J., 128, 1748-1760 (2004) Photometry of type II Cepheids. I. The long-period stars.

J/AJ/128/1761/  lc/* table1.dat  (51.0 Mb)

Hartman J.D., Bakos G., Stanek K.Z., Noyes R.W. Astron. J., 128, 1761-1783 (2004) HATNET variability survey in the high stellar density ”Kepler field” with millimagnitude image subtraction photometry.

J/AJ/128/1825/  table1.dat table2.dat table3.dat  (17Kb)

Kilic M., Winget D.E., von Hippel T., Claver C.F. Astron. J., 128, 1825-1833 (2004) Photometric identification of cool white dwarfs.

J/AJ/128/1840/  table3.dat  (160Kb)

Lacy C.H.S., Claret A., Sabby J.A. Astron. J., 128, 1840-1846 (2004) Absolute properties of the upper main-sequence eclipsing binary star MU Cassiopeiae.

J/AJ/128/1974/  table2.dat  (29Kb)

Schinnerer E., Carilli C.L., Scoville N.Z., Bondi M., Ciliegi P., Vettolani P., Le Fevre O., Koekemoer A.M., Bertoldi F., Impey C.D. Astron. J., 128, 1974-1989 (2004) The VLA-COSMOS survey. I. Radio identifications from the pilot project.

J/AJ/128/1990/  table2.dat  (12Kb)

Cross N.J.G., Bouwens R.J., Benitez N., Blakeslee J.P., Menanteau F., Ford H.C., Goto T., Holden B., Martel A.R., Zirm A., Overzier R., Gronwall C., Homeier N., Clampin M., Hartig G.F., Illingworth G.D., Ardila D.R., Bartko F., Broadhurst T.J., Brown R.A., Burrows C.J., Cheng E.S., Feldman P.D., Franx M., Golimowski D.A., Infante L., Kimble R.A., Krist J.E., Lesser M.P., Meurer G.R., Miley G.K., Postman M., Rosati P., Sirianni M., Sparks W.B., Tran H.D., Tsvetanov Z.I., White R.L., Zheng W. Astron. J., 128, 1990-2012 (2004) The luminosity function of early-type field galaxies at z 0.75.

J/AJ/128/2080/  refs.dat table1.dat table2.dat table3.dat  (40Kb)

O’Neil K. Astron. J., 128, 2080-2097 (2004) New reference galaxy standards for H I emission observations.

J/AJ/128/2170/  refs.dat table3.dat table4.dat  (27Kb)

Hunter D.A., Elmegreen B.G. Astron. J., 128, 2170-2205 (2004)

Star formation properties of a large sample of irregular galaxies.

J/AJ/128/2239/  table2.dat table3.dat  (144Kb)

Persson S.E., Madore B.F., Krzeminski W., Freedman W.L., Roth M., Murphy D.C. Astron. J., 128, 2239-2264 (2004) New cepheid period-luminosity relations for the Large Magellanic Cloud: 92 near-infrared light curves.

J/AJ/128/2316/  table3.dat table4.dat  (38Kb)

Sherry W.H., Walter F.M., Wolk S.J. Astron. J., 128, 2316-2330 (2004) Photometric identification of the low-mass population of Orion OB1b. I. The σOrionis cluster.

J/AJ/128/2364/  table1.dat  (17Kb)

Roman-lopes A., Abraham Z. Astron. J., 128, 2364-2373 (2004) A rich star-forming region in the direction of RCW 95.

J/AJ/128/2402/  ew.dat refs.dat stars.dat table2.dat table7.dat table8.dat table9.dat  (86Kb)

Lai D.K., Bolte M., Johnson J.A., Lucatello S. Astron. J., 128, 2402-2419 (2004) Abundances of extremely metal-poor star candidates.

J/AJ/128/2642/  refs.dat table1.dat table5.dat table6.dat  (90Kb)

Caretta C.A., Maia M.A.G., Willmer C.N.A. Astron. J., 128, 2642-2651 (2004) The Aquarius superclusters. II. Spectroscopic and photometric data.

J/AJ/128/2815/  table1.dat table2.dat  (480Kb)

Pietrzynski G., Gieren W., Udalski A., Bresolin F., Kudritzki R.-P., Soszynski I., Szymanski M., Kubiak M. Astron. J., 128, 2815-2825 (2004) The Araucaria project: the distance to the Local Group galaxy NGC 6822 from cepheid variables discovered in a wide-field imaging survey.

J/AJ/128/2965/  table4.dat  (988Kb)

Williams P.R., Wozniak S.J., Vestrand W.T., Gupta V. Astron. J., 128, 2965-2976 (2004) Identifying red variables in the Northern Sky Variability Survey.

J/AJ/128/3012/  table1.dat table2.dat table3.dat table4.dat table5.dat table6.dat  (104Kb)

Mason B.D., Hartkopf W.I., Wycoff G.L., Rafferty T.J., Urban S.E., Flagg L. Astron. J., 128, 3012-3018 (2004) Speckle interferometry at the US Naval Observatory. X.

J/AJ/128/3019/  table1.dat  (33Kb)

Clark L.L., Sandquist E.L., Bolte M. Astron. J., 128, 3019-3033 (2004) The blue straggler and main-sequence binary population of the low-mass globular cluster Palomar 13.

J/AJ/128/3053/  table3.dat table4.dat  (668Kb)

Bohlin R.C., Gilliland R.L. Astron. J., 128, 3053-3060 (2004) Absolute flux distribution of the SDSS standard BD +17^{∘}4708.

J/AJ/129/26/  table1.dat  (36Kb)

Owen F.N., Keel W.C., Ledlow M.J., Morrison G.E., Windhorst R.A. Astron. J., 129, 26-30 (2005) A deep radio survey of Abell 2125. I. Radio, optical, and near-infrared observations.

J/AJ/129/31/  table1.dat table2.dat table4.dat table8.dat  (33Kb)

Owen F.N., Ledlow M.J., Keel W.C., Wang Q.D., Morrison G.E. Astron. J., 129, 31-52 (2005) A deep radio survey of Abell 2125. II. Accelerated galaxy evolution during a cluster-cluster merger.

J/AJ/129/220/  table1.dat table2.dat  (22Kb)

Donley J.L., Staveley-Smith L., Kraan-Korteweg R.C., Islas-Islas J.M., Schroeder A., Henning P.A., Koribalski B., Mader S., Stewart I. Astron. J., 129, 220-238 (2005) The H I Parkes zone of avoidance survey: the northern extension.

J/AJ/129/267/ table1.dat table2.dat table3.dat table6.dat table7.dat table8.dat table9.dat table10.dat table11.dat table12.dat table13.dat (65Kb) Cacciari C., Corwin T.M., Carney B.W.

Astron. J., 129, 267-302 (2005) A multicolor and Fourier study of RR Lyrae variables in the globular cluster NGC 5272 (M3).

J/AJ/129/303/  table1a.dat table1b.dat table2a.dat table2b.dat table2c.dat  (483Kb)

Cohen J.G., Melendez J. Astron. J., 129, 303-329 (2005) Abundances in a large sample of stars in M3 and M13.

J/AJ/129/348/  table4a.dat table4b.dat  (120Kb)

Giveon U., Becker R.H., Helfand D.J., White R.L. Astron. J., 129, 348-354 (2005) A new catalog of radio compact H II regions in the Milky Way.

J/AJ/129/363/  table1.dat table2.dat  (32Kb)

Sicilia-Aguilar A., Hartmann L.W., Szentgyorgyi A.H., Fabricant D.G., Furesz G., Roll J., Conroy M.A., Calvet N., Tokarz S., Hernandez J. Astron. J., 129, 363-381 (2005) Accretion, kinematics, and rotation in the Orion Nebula Cluster: initial results from Hectochelle.

J/AJ/129/393/  table2.dat  (36Kb)

Oey M.S., Watson A.M., Kern K., Walth G.L. Astron. J., 129, 393-401 (2005) Hierarchical triggering of star formation by superbubbles in W3/W4.

J/AJ/129/413/  table2.dat  (24Kb)

Subasavage J.P., Henry T.J., Hambly N.C., Brown M.A., Jao W.-C. Astron. J., 129, 413-432 (2005) The solar neighborhood. XII. Discovery of new high proper motion stars with μ=0.4/yr between declinations -90^{∘}and -47^{∘}.

J/AJ/129/433/  table5.dat table6.dat table7.dat  (6.6 Mb)

Kovtyukh V.V., Andrievsky S.M., Belik S.I., Luck R.E. Astron. J., 129, 433-453 (2005) Phase-dependent variation of the fundamental parameters of cepheids. II. Periods longer than 10 days.

J/AJ/129/656/  table1.dat  (169Kb)

Carney B.W., Lee J.-W., Dodson B. Astron. J., 129, 656-668 (2005) Infrared photometry of NGC 6791.

J/AJ/129/682/  table1.dat table2.dat  (17Kb)

Hernandez-Toledo H.M., Avila-Reese V., Conselice C.J., Puerari I. Astron. J., 129, 682-697 (2005) The structural properties of isolated galaxies, spiral-spiral pairs, and mergers: the robustness of galaxy morphology during secular evolution.

J/AJ/129/729/  table4.dat  (1.3 Mb)

Rowe J.F., Richer H.B., Brewer J.P., Crabtree D.R. Astron. J., 129, 729-744 (2005) Carbon stars and other luminous stellar populations in M33.

J/AJ/129/809/  table1.dat table2.dat table3.dat  (47Kb)

Strom S.E., Wolff S.C., Dror D.H.A. Astron. J., 129, 809-828 (2005) B star rotational velocities in h and χPersei: a probe of initial conditions during the star formation epoch.

J/AJ/129/829/  table1.dat  (63Kb)

Dahm S.E., Simon T. Astron. J., 129, 829-855 (2005) The T Tauri star population of the young cluster NGC 2264.

J/AJ/129/856/  table2.dat table3.dat  (47Kb)

Hernandez J., Calvet N., Hartmann L., Briceno C., Sicilia-Aguilar A., Berlind P. Astron. J., 129, 856-871 (2005) Herbig Ae/Be stars in nearby OB associations.

J/AJ/129/872/  table1.dat  (152Kb)

Anthony-Twarog B.J., Atwell J., Twarog B.A. Astron. J., 129, 872-887 (2005) uvbyCaHβCCD photometry of clusters. V. The metal-deficient open cluster NGC 2243.

J/AJ/129/907/  table3.dat table4.dat table6.dat table7.dat  (45Kb)

Briceno C., Calvet N., Hernandez J., Vivas A.K., Hartmann L., Downes J.J., Berlind P. Astron. J., 129, 907-926 (2005) The CIDA variability survey of Orion OB1. I. The low-mass population of Ori OB1a and 1b.

J/AJ/129/1001/  table1.dat  (10Kb)

Fekel F.C., Barlow D.J., Scarfe C.D., Jancart S., Pourbaix D. Astron. J., 129, 1001-1007 (2005) HD 166181 = V815 Herculis, a single-lined spectroscopic multiple system.

J/AJ/129/1049/  table1.dat  (12Kb)

Carpenter J.M., Wolf S., Schreyer K., Launhardt R., Henning T. Astron. J., 129, 1049-1062 (2005) Evolution of cold circumstellar dust around solar-type stars.

J/AJ/129/1063/  table1.dat table2.dat table3.dat table4.dat table5.dat table7.dat  (226Kb)

Luck R.E., Heiter U. Astron. J., 129, 1063-1083 (2005) Stars within 15 parsecs: abundances for a northern sample.

J/AJ/129/1163/  table2.dat  (46Kb)

Petrov L., Kovalev Y.Y., Fomalont E., Gordon D. Astron. J., 129, 1163-1170 (2005) The third VLBA calibrator survey: VCS3.

J/AJ/129/1198/  table1.dat table2.dat table3.dat table4.dat  (41Kb)

Hatziminaoglou E., Perez-Fournon I., Polletta M., Afonso-Luis A., Hernan-Caballero A., Montenegro-Montes F.M., Lonsdale C., Xu C.K., Franceschini A., Rowan-Robinson M., Babbedge T., Smith H.E., Surace J., Shupe D., Fang F., Farrah D., Oliver S., Gonzalez-Solares E.A., Serjeant S. Astron. J., 129, 1198-1211 (2005) Sloan Digital Sky Survey quasars in the Spitzer wide-area infrared extragalactic survey (SWIRE) ELAIS N1 field: properties and spectral energy distributions.

J/AJ/129/1237/  table1.dat table3.dat  (64Kb)

Coenda V., Donzelli C.J., Muriel H., Quintana H., Infante L., Lambas D.G. Astron. J., 129, 1237-1248 (2005) Photometric observations of Southern Abell Clusters Redshifts Survey clusters: galaxy luminosity profiles.

J/AJ/129/1311/  table1.dat table2.dat  (40Kb)

Rosenberg J.L., Schneider S.E., Posson-Brown J. Astron. J., 129, 1311-1330 (2005) Gas and stars in an H I-selected galaxy sample.

J/AJ/129/1465/  table2.dat table3.dat  (50Kb)

Cole A.A., Tolstoy E., Gallagher J.S.III, Smecker-Hane T.A. Astron. J., 129, 1465-1482 (2005) Spectroscopy of red giants in the Large Magellanic Cloud bar: abundances, kinematics, and the age-metallicity relation.

J/AJ/129/1483 See I/298/

J/AJ/129/1534/  table3.dat table4.dat table5.dat  (39Kb)

Robberto M., Beckwith S.V.W., Panagia N., Patel S.G., Herbst T.M., Ligori S., Custo A., Boccacci P., Bertero M. Astron. J., 129, 1534-1563 (2005) The Orion nebula in the mid-infrared.

J/AJ/129/1596/  table23.dat  (32Kb)

Hartman J.D., Kaluzny J., Szentgyorgyi A., Stanek K.Z. Astron. J., 129, 1596-1606 (2005) BVI photometric variability survey of M3.

J/AJ/129/1607/  table3.dat  (15Kb)

Cohen J.G., Melendez J. Astron. J., 129, 1607-1615 (2005) Outer versus inner halo globular clusters: NGC 7492 abundances.

J/AJ/129/1642/  table8.dat table9.dat table10.dat table11.dat  (2.0 Mb)

Fitzpatrick E.L., Massa D. Astron. J., 129, 1642-1662 (2005) Determining the physical properties of the B stars. II. Calibration of synthetic photometry.

J/AJ/129/1669/  table4.dat table5.dat table6.dat  (159Kb)

Fekel F.C., Henry G.W. Astron. J., 129, 1669-1685 (2005) Chromospherically active stars. XXIV. The giant, single-lined binaries HD 37824, HD 181809, and HD 217188.

J/AJ/129/1755 See II/259/

J/AJ/129/1818/  table1.dat table2.dat table3.dat table4.dat table5.dat table6.dat  (45Kb)

Wu J., Peng B., Zhou X., Ma J., Jiang Z., Chen J. Astron. J., 129, 1818-1826 (2005) Optical monitoring of BL Lacertae object S5 0716+714 with high temporal resolution.

J/AJ/129/1832/  table6.dat table7.dat table8.dat  (16Kb)

Grutzbauch R., Kelm B., Focardi P., Trinchieri G., Rampazzo R., Zeilinger W.W. Astron. J., 129, 1832-1848 (2005) Small-scale systems of galaxies. II. Properties of the NGC 4756 group of galaxies.

J/AJ/129/2007/  table1.dat table2.dat table3.dat table4.dat  (45Kb)

Schmidt E.G., Johnston D., Langan S., Lee K.M. Astron. J., 129, 2007-2017 (2005) Photometry of type II Cepheids. II. The short-period stars.

J/AJ/129/2018/  table2.dat  (69Kb)

Corcoran M.F. Astron. J., 129, 2018-2025 (2005) X-ray monitoring of ηCarinae: variations on a theme.

J/AJ/129/2026/  table1.dat  (26Kb)

Henry G.W., Fekel F.C. Astron. J., 129, 2026-2033 (2005) HD 8801: a unique single Am star with γDoradus and δScuti pulsations.

J/AJ/129/2062/  table1.dat  (199Kb)

Allam S.S., Tucker D.L., Lee B.C., Smith J.A. Astron. J., 129, 2062-2073 (2005) A catalog of very isolated galaxies from the Sloan Digital Sky Survey data release 1.

J/AJ/129/2074/  table2.dat  (19Kb)

Franceschini A., Manners J., Polletta M.D.C., Lonsdale C., Gonzalez-Solares E., Surace J., Shupe D., Fang F., Xu C.K., Farrah D., Berta S., Rodighiero G., Perez-Fournon I., Hatziminaoglou E., Smith H.E., Siana B., Rowan-Robinson M., Nandra K., Babbedge T., Vaccari M., Oliver S., Wilkes B., Owen F., Padgett D., Frayer D., Jarrett T., Masci F., Stacey G., Almaini O., McMahon R., Johnson O., Lawrence A., Willott C. Astron. J., 129, 2074-2101 (2005) A complete multiwavelength characterization of faint Chandra X-ray sources seen in the Spitzer wide-area infrared extragalactic (SWIRE) survey.

J/AJ/129/2119/  table1.dat  (19Kb)

Spekkens K., Giovanelli R., Haynes M.P. Astron. J., 129, 2119-2137 (2005) The cusp/core problem in galactic halos: long-slit spectra for a large dwarf galaxy sample.

J/AJ/129/2138/  table2.dat table12.dat  (145Kb)

Lauer T.R., Faber S.M., Gebhardt K., Richstone D., Tremaine S., Ajhar E.A., Aller M.C., Bender R., Dressler A., Filippenko A.V., Green R., Grillmair C.J., Ho L.C., Kormendy J., Magorrian J., Pinkney J., Siopis C. Astron. J., 129, 2138-2185 (2005) The centers of early-type galaxies with Hubble Space Telescope. V. New WFPC2 photometry.

J/AJ/129/2232/  table2.dat table3.dat table4.dat table5.dat table6.dat table7.dat  (151Kb)

Pritzl B.J., Armandroff T.E., Jacoby G.H., Da Costa G.S. Astron. J., 129, 2232-2256 (2005) The dwarf spheroidal companions to M31: variable stars in Andromeda I and Andromeda III.

J/AJ/129/2294/  table1.dat table5.dat  (27Kb)

Massarotti A., Latham D.W., Torres G., Brown R.A., Oppenheimer B.D. Astron. J., 129, 2294-2307 (2005) A search for substellar companions around 15 weak-lined T Tauri stars with the Planetary Camera 2 of the Hubble Space Telescope.

J/AJ/129/2420/  table1.dat table2.dat  (483Kb)

Makarov V.V., Kaplan G.H. Astron. J., 129, 2420-2427 (2005) Statistical constraints for astrometric binaries with nonlinear motion.

J/AJ/129/2428/  table1.dat table2.dat table3.dat table4.dat table5.dat  (70Kb)

Silvestri N.M., Hawley S.L., Oswalt T.D. Astron. J., 129, 2428-2450 (2005) The chromospheric activity and ages of M dwarf stars in wide binary systems.

J/AJ/129/2542/  table1.dat table2.dat table3.dat table4.dat table5.dat  (20Kb)

Collinge M.J., Strauss M.A., Hall P.B., Ivezic Z., Munn J.A., Schlegel D.J., Zakamska N.L., Anderson S.F., Harris H.C., Richards G.T., Schneider D.P., Voges W., York D.G., Margon B., Brinkmann J. Astron. J., 129, 2542-2561 (2005) Optically identified BL Lacertae objects from the Sloan Digital Sky Survey.

J/AJ/129/2663/  table1.dat  (24Kb)

Williams B.F. Astron. J., 129, 2663-2669 (2005) Variable star candidates in an advanced camera for surveys field of M31.

J/AJ/129/2701/  table1.dat table2.dat  (89Kb)

Rafelski M., Zaritsky D. Astron. J., 129, 2701-2713 (2005) The star clusters of the Small Magellanic Cloud: age distribution.

J/AJ/129/2714/  stars.dat table1.dat  (26Kb)

Del Principe M., Piersimoni A.M., Bono G., Di Paola A., Dolci M., Marconi M. Astron. J., 129, 2714-2724 (2005) Near-infrared observations of RR Lyrae variables in galactic globular clusters. I. The case of M92.

J/AJ/129/2725/  table3.dat  (8Kb)

Friel E.D., Jacobson H.R., Pilachowski C.A. Astron. J., 129, 2725-2730 (2005) Abundances of red giants in old open clusters. II. Berkeley 17.

J/AJ/129/2798/  table1.dat table2.dat  (24Kb)

Williamon R.M., Van Hamme W., Torres G., Sowell J.R., Ponce V.C. Astron. J., 129, 2798-2805 (2005) The eclipsing binary system AR Monocerotis.

J/AJ/129/2815/  table3.dat table4.dat table5.dat  (209Kb)

Henry G.W., Fekel F.C., Henry S.M. Astron. J., 129, 2815-2830 (2005) Eleven new γDoradus stars.

J/AJ/129/2856/  table4.dat table5.dat table6.dat  (13Kb)

Mochejska B.J., Stanek K.Z., Sasselov D.D., Szentgyorgyi A.H., Bakos G.A., Hradecky V., Devor J., Marrone D.P., Winn J.N., Zaldarriaga M. Astron. J., 129, 2856-2868 (2005) Planets in stellar clusters extensive search. III. A search for transiting planets in the metal-rich open cluster NGC 6791.

J/AJ/130/1/  table8.dat  (314Kb)

Thompson R.I., Illingworth G., Bouwens R., Dickinson M., Eisenstein D., Fan X., Franx M., Riess A., Rieke M.J., Schneider G., Stobie E., Toft S., Van Dokkum P. Astron. J., 130, 1-12 (2005) The near-infrared camera and multi-object spectrometer ultra deep field: observations, data reduction, and galaxy photometry.

J/AJ/130/23/  notes.dat table1.dat table3.dat  (31Kb)

Aars C.E., Hough D.H., Yu L.H., Linick J.P., Beyer P.J., Vermeulen R.C., Readhead A.C.S. Astron. J., 130, 23-46 (2005) Optical spectrophotometry of a complete sample of 3CR lobe-dominated quasars.

J/AJ/130/47/  table1.dat table2.dat  (22Kb)

Ledlow M.J., Owen F.N., Miller N.A. Astron. J., 130, 47-54 (2005) The cluster of galaxies surrounding Cygnus A. II. New velocities and a dynamic model.

J/AJ/130/55/  table8.dat  (17Kb)

Mendes De Oliveira C., Coelho P., Gonzalez J.J., Barbuy B. Astron. J., 130, 55-64 (2005) Ages, metallicities, and α-element enhancement for galaxies in Hickson compact groups.

J/AJ/130/156/  table4a.dat table4b.dat  (108Kb)

Giveon U., Becker R.H., Helfand D.J., White R.L. Astron. J., 130, 156-164 (2005) A new catalog of radio compact H II regions in the Milky way. II. The 1.4 GHz data.

J/AJ/130/188/ table1.dat table2.dat table3.dat table4.dat table5.dat table6.dat table7.dat table8.dat (72Kb) Sicilia-Aguilar A., Hartmann L.W., Hernandez J., Briceno C., Calvet N. Astron. J., 130, 188-209 (2005) Cepheus OB2: disk evolution and accretion at 3-10 Myr.

J/AJ/130/234/  table3.dat table4.dat table5.dat table6.dat table7.dat  (23Kb)

Mader J.A., Torres G., Marschall L.A., Rizvi A. Astron. J., 130, 234-245 (2005) Optical photometry and spectroscopy of the suspected “Cool Algol” AV Delphini: determination of the physical properties.

J/AJ/130/367 See VII/243/

J/AJ/130/381/  table1.dat  (10Kb)

Boroson T. Astron. J., 130, 381-386 (2005) Blueshifted [O III] emission: indications of a dynamic narrow-line region.

J/AJ/130/387/  table1.dat  (23Kb)

Strateva I.V., Brandt W.N., Schneider D.P., Vanden Berk D.E., Vignali C. Astron. J., 130, 387-405 (2005) Soft X-ray and ultraviolet emission relations in optically selected AGN samples.

J/AJ/130/425/  table1.dat table2.dat  (172Kb)

De Carvalho R.R., Goncalves T.S., Iovino A., Kohl-Moreira J.L., Gal R.R., Djorgovski S.G. Astron. J., 130, 425-444 (2005) A catalog of distant compact groups using the Digitized Second Palomar Observatory Sky Survey.

J/AJ/130/445/  table7.dat table8.dat table9.dat  (21Kb)

Barr J., Davies R., Jorgensen I., Bergmann M., Crampton D. Astron. J., 130, 445-474 (2005) RX J0142.0+2131. I. The galaxy content of an X-ray-luminous galaxy cluster at z=0.28.

J/AJ/130/496/  table1.dat table2.dat table3.dat table4.dat  (47Kb)

Jangren A., Wegner G., Salzer J.J., Werk J.K., Gronwall C. Astron. J., 130, 496-505 (2005) Spectroscopy of KISS emission-line galaxy candidates. III. A second set of MDM observations.

J/AJ/130/586/  table3.dat table4.dat table5.dat table6.dat  (1.2 Mb)

White R.L., Becker R.H., Helfand D.J. Astron. J., 130, 586-596 (2005) New catalogs of compact radio sources in the galactic plane.

J/AJ/130/652/  table1.dat table3.dat  (9Kb)

Villanova S., Carraro G., Bresolin F., Patat F. Astron. J., 130, 652-658 (2005) Metal abundances in extremely distant galactic old open clusters. II. Berkeley 22 and Berkeley 66.

J/AJ/130/767/  stars.dat table1.dat  (34Kb)

Rucinski S.M., Pych W., Ogloza W., Debond H., Thomson J.R., Mochnacki S.W., Capobianco C.C., Conidis G., Rogoziecki P. Astron. J., 130, 767-775 (2005) Radial velocity studies of close binary stars. X.

J/AJ/130/776/  table1.dat  (32Kb)

Templeton M.R., Mattei J.A., Willson L.A. Astron. J., 130, 776-788 (2005) Secular evolution in Mira variable pulsations.

J/AJ/130/794/  table1.dat table2.dat  (23Kb)

Fekel F.C., Henry G.W., Lewis C. Astron. J., 130, 794-798 (2005) Chromospherically active stars. XXV. HD 144110=EV Draconis, a double-lined dwarf binary.

J/AJ/130/825/  table2.dat  (18Kb)

Welsh B.Y., Wheatley J.M., Heafield K., Seibert M., Browne S.E., Salim S., Rich R.M., Barlow T.A., Bianchi L., Byun Y.-I., Donas J., Forster K., Friedman P.G., Heckman T.M., Jelinsky P.N., Lee Y.-W., Madore B.F., Malina R.F., Martin D.C., Milliard B., Morrissey P., Neff S.G., Schiminovich D., Siegmund O.H.W., Small T., Szalay A.S., Wyder T.K. Astron. J., 130, 825-831 (2005) The GALEX ultraviolet variability catalog.

J/AJ/130/832/  table2.dat table3.dat  (39Kb)

Schmidt E.G., Johnston D., Langan S., Lee K.M. Astron. J., 130, 832-841 (2005) Photometry of type II cepheids. III. The intermediate-period stars.

J/AJ/130/873/  table2.dat  (12Kb)

Jester S., Schneider D.P., Richards G.T., Green R.F., Schmidt M., Hall P.B., Strauss M.A., Vanden Berk D.E., Stoughton C., Gunn J.E., Brinkmann J., Kent S.M., Smith J.A., Tucker D.L., Yanny B. Astron. J., 130, 873-895 (2005) The Sloan Digital Sky Survey view of the Palomar-Green Bright Quasar Survey.

J/AJ/130/896/  table5.dat table6.dat table7.dat table8.dat tablea2.dat tablea31.dat  (24Kb)

Saripalli L., Hunstead R.W., Subrahmanyan R., Boyce E. Astron. J., 130, 896-922 (2005) A complete sample of megaparsec-sized double radio sources from the Sydney University Molonglo Sky Survey.

J/AJ/130/923/  notes.dat table1.dat  (32Kb)

Wrobel J.M., Taylor G.B., Rector T.A., Myers S.T., Fassnacht C.D. Astron. J., 130, 923-935 (2005) Faint radio sources in the NOAO Bootes field: VLBA imaging and optical identifications.

J/AJ/130/957/  table2.dat  (38Kb)

Fleenor M.C., Rose J.A., Christiansen W.A., Hunstead R.W., Johnston-Hollitt M., Drinkwater M.J., Saunders W. Astron. J., 130, 957-967 (2005) Large-scale velocity structures in the Horologium-Reticulum supercluster.

J/AJ/130/968/  table2.dat  (114Kb)

Miller C.J., Nichol R.C., Reichart D., Wechsler R.H., Evrard A.E., Annis J., McKay T.A., Bahcall N.A., Bernardi M., Boehringer H., Connolly A.J., Goto T., Kniazev A., Lamb D., Postman M., Schneider D.P., Sheth R.K., Voges W. Astron. J., 130, 968-1001 (2005) The C4 clustering algorithm: clusters of galaxies in the Sloan Digital Sky Survey.

J/AJ/130/1002/  a2125.dat a2218.dat  (44Kb)

Rakos K., Schombert J. Astron. J., 130, 1002-1021 (2005) Cluster populations in Abell 2125 and 2218.

J/AJ/130/1037/  table1.dat  (47Kb)

Catinella B., Haynes M.P., Giovanelli R. Astron. J., 130, 1037-1048 (2005) Rotational widths for use in the Tully-Fisher relation. I. Long-slit spectroscopic data.

J/AJ/130/1065/  table7.dat  (17Kb)

Coleman M.G., Da Costa G.S., Bland-hawthorn J. Astron. J., 130, 1065-1082 (2005) The absence of extratidal structure in the Sculptor dwarf spheroidal galaxy.

J/AJ/130/1083/  table4.dat table5.dat table6.dat table7.dat table8.dat  (936Kb)

McCumber M.P., Garnett D.R., Dufour R.J. Astron. J., 130, 1083-1096 (2005) Stellar populations in the wing of the Small Magellanic Cloud from Hubble Space Telescope photometry.

J/AJ/130/1097/  table2.dat table3.dat  (43Kb)

Brown W.R., Geller M.J., Kenyon S.J., Kurtz M.J., Allende Prieto C., Beers T.C., Wilhelm R. Astron. J., 130, 1097-1110 (2005) The Century Survey Galactic Halo Project. II. Global properties and the luminosity function of field blue horizontal branch stars.

J/AJ/130/1111/  table2.dat table3.dat table6.dat  (424Kb)

Carney B.W., Yong D., De Almeida M.L.T., Seitzer P. Astron. J., 130, 1111-1126 (2005) Elemental abundance ratios in stars of the outer galactic disk. II. Field red giants.

J/AJ/130/1145/  table1.dat table3.dat  (21Kb)

Doppmann G.W., Greene T.P., Covey K.R., Lada C.J. Astron. J., 130, 1145-1170 (2005) The physical natures of class I and flat-spectrum protostellar photospheres: a near-infrared spectroscopic study.

J/AJ/130/1177/  table1.dat table2.dat table3.dat table4.dat  (25Kb)

Cohen J.G., Briley M.M., Stetson P.B. Astron. J., 130, 1177-1193 (2005) C and N abundances in stars at the base of the red giant branch in M15.

J/AJ/130/1194/  refs.dat table1.dat table2.dat table5.dat table6.dat  (20Kb)

Turner D.G., Forbes D., Van Den Bergh S., Younger P.F., Berdnikov L.N. Astron. J., 130, 1194-1205 (2005) The open cluster Ruprecht 91 and its Cepheids.

J/AJ/130/1247/  table1.dat  (25Kb)

Lepine S. Astron. J., 130, 1247-1260 (2005) New high proper motion stars from the digitized sky survey. III. Stars with proper motions 0.45” μ2.0”/yr south of declination -30^{∘}.

J/AJ/130/1324/  table2.dat  (149Kb)

Drozdovsky I., Yan L., Chen H.-W., Stern D., Kennicutt R.JR, Spinrad H., Dawson S. Astron. J., 130, 1324-1336 (2005) The Hubble Space Telescope ACS grism parallel survey. II. First results and a catalog of faint emission-line galaxies at z =1.6.

J/AJ/130/1345/  table2.dat table3.dat  (13Kb)

Ellison S.L., Hall P.B., Lira P. Astron. J., 130, 1345-1357 (2005) The optical-infrared colors of CORALS QSOs: searching for dust reddening associated with high-redshift damped Lyαsystems.

J/AJ/130/1358/  table3.dat table4.dat  (23Kb)

Norris R.P., Huynh M.T., Jackson C.A., Boyle B.J., Ekers R.D., Mitchell D.A., Sault R.J., Wieringa M.H., Williams R.E., Hopkins A.M., Higdon J. Astron. J., 130, 1358-1372 (2005) Radio observations of the Hubble Deep Field-South region. I. Survey description and initial results.

J/AJ/130/1373/  table1.dat  (55Kb)

Huynh M.T., Jackson C.A., Norris R.P., Prandoni I. Astron. J., 130, 1373-1388 (2005) Radio observations of the Hubble Deep Field-South region. II. The 1.4 GHz catalog and source counts.

J/AJ/130/1389/  table1.dat table3.dat table4.dat table5.dat  (61Kb)

Lister M.L., Homan D.C. Astron. J., 130, 1389-1417 (2005) MOJAVE: Monitoring of Jets in Active Galactic Nuclei with VLBA Experiments. I. First-epoch 15GHz linear polarization images.

J/AJ/130/1418/  table1.dat table5.dat  (20Kb)

Jorstad S.G., Marscher A.P., Lister M.L., Stirling A.M., Cawthorne T.V., Gear W.K., Gomez J.L., Stevens J.A., Smith P.S., Forster J.R., Robson E.I. Astron. J., 130, 1418-1465 (2005) Polarimetric observations of 15 active galactic nuclei at high frequencies: jet kinematics from bimonthly monitoring with the Very Long Baseline Array.

J/AJ/130/1502/  table1.dat  (38Kb)

Mahdavi A., Trentham N., Tully R.B. Astron. J., 130, 1502-1515 (2005) The NGC 5846 group: dynamics and the luminosity function to m_{R}=-12.

J/AJ/130/1516/  table1.dat  (8Kb)

De Propris R., Liske J., Driver S.P., Allen P.D., Cross N.J.G. Astron. J., 130, 1516-1523 (2005) The Millennium Galaxy Catalogue: dynamically close pairs of galaxies and the global merger rate.

J/AJ/130/1640/  table1.dat  (10Kb)

Wu C., Qiu Y.L., Deng J.S., Hu J.Y., Zhao Y.H. Astron. J., 130, 1640-1651 (2005) CCD photometry of candidate RR Lyrae stars in the Sloan Digital Sky Survey.

J/AJ/130/1652 See V/125/

J/AJ/130/1658/  table2.dat  (26Kb)

Subasavage J.P., Henry T.J., Hambly N.C., Brown M.A., Jao W.-C., Finch C.T. Astron. J., 130, 1658-1679 (2005) The solar neighborhood. XV. Discovery of new high proper motion stars with μ=0.4”/yr between declinations -47^{∘}and 00^{∘}.

J/AJ/130/1680/  refs.dat table2.dat table3.dat table4.dat  (509Kb)

Lepine S. Astron. J., 130, 1680-1692 (2005) Nearby stars from the LSPM-North proper-motion catalog. I. Main-sequence dwarfs and giants within 33 parsecs of the Sun.

J/AJ/130/1693/ table2.dat table3.dat table4.dat table5.dat table6.dat table7.dat table8.dat table9.dat table10.dat table11.dat table12.dat table13.dat (74Kb) Brown T.M., Ferguson H.C., Smith E., Guhathakurta P., Kimble R.A., Sweigart A.V., Renzini A., Rich R.M., Vandenberg D.A. Astron. J., 130, 1693-1706 (2005) Stellar cluster fiducial sequences with the Advanced Camera for Surveys.

J/AJ/130/1707/  table1.dat  (11Kb)

Vasconcelos M.J., Cerqueira A.H., Plana H., Raga A.C., Morisset C. Astron. J., 130, 1707-1718 (2005) Gemini multi-object spectrograph integral field unit spectroscopy of the 167-317 (LV2) proplyd in Orion.

J/AJ/130/1719/  table3.dat  (12Kb)

Leistra A., Cotera A.S., Liebert J., Burton M. Astron. J., 130, 1719-1732 (2005) Two 2MASS-selected young stellar clusters: photometry, spectroscopy, and the initial mass function.

J/AJ/130/1733/  table2.dat table3.dat table4.dat  (45Kb)

Wilking B.A., Meyer M.R., Robinson J.G., Greene T.P. Astron. J., 130, 1733-1751 (2005) Optical spectroscopy of the surface population of the ρOphiuchi molecular cloud: the first wave of star formation.

J/AJ/130/1763/  table1.dat  (12Kb)

Smith N., Bally J., Shuping R.Y., Morris M., Kassis M. Astron. J., 130, 1763-1777 (2005) Thermal dust emission from proplyds, unresolved disks, and shocks in the Orion nebula.

J/AJ/130/1805/  table2.dat  (16Kb)

Dahm S.E. Astron. J., 130, 1805-1828 (2005) The young cluster NGC 2362.

J/AJ/130/1871/  table1.dat table2.dat table3.dat  (79Kb)

Bochanski J.J., Hawley S.L., Reid I.N., Covey K.R., West A.A., Tinney C.G., Gizis J.E. Astron. J., 130, 1871-1879 (2005) Spectroscopic survey of M dwarfs within 100 parsecs of the Sun.

J/AJ/130/1880/  tablea1.dat  (87Kb)

Andrievsky S.M., Luck R.E., Kovtyukh V.V. Astron. J., 130, 1880-1889 (2005) Phase-dependent variation of the fundamental parameters of cepheids. III. Periods between 3 and 6 days.

J/AJ/130/1977/  table1.dat table2.dat  (16Kb)

Haarsma D.B., Winn J.N., Falco E.E., Kochanek C.S., Ammar P., Boersma C., Fogwell S., Muxlow T.W.B., Mcleod B.A., Lehar J. Astron. J., 130, 1977-1995 (2005) The FIRST-Optical-VLA survey for lensed radio lobes.

J/AJ/130/2012/  table2.dat  (193Kb)

Way M.J., Quintana H., Infante L., Lambas D.G., Muriel H. Astron. J., 130, 2012-2018 (2005) Redshifts in the Southern Abell Redshift Survey clusters. I. The data.

J/AJ/130/2043/  table2.dat table3.dat  (17Kb)

Patton D.R., Grant J.K., Simard L., Pritchet C.J., Carlberg R.G., Borne K.D. Astron. J., 130, 2043-2057 (2005)

A Hubble Space Telescope snapshot survey of dynamically close galaxy pairs in the CNOC2 redshift survey.

J/AJ/130/2140/  table2.dat table3.dat  (36Kb)

Pritzl B.J., Venn K.A., Irwin M. Astron. J., 130, 2140-2165 (2005) A comparison of elemental abundance ratios in globular clusters, field stars, and dwarf spheroidal galaxies.

J/AJ/130/2212/  table1.dat table3.dat  (25Kb)

Tsujimoto M., Kobayashi N., Tsuboi Y. Astron. J., 130, 2212-2219 (2005) X-ray and near-infrared studies of the star-forming cloud L1448.

J/AJ/130/2252/  table2.dat table3.dat table4.dat  (14Kb)

Yang Y.-G., Qian S.-B., Zhu L.-Y. Astron. J., 130, 2252-2261 (2005) Deep, low mass ratio overcontact binary systems. IV. V410 Aurigae and XY Bootis.

J/AJ/130/2439/  table5.dat  (6.5 Mb)

Suchkov A.A., Hanisch R.J., Margon B. Astron. J., 130, 2439-2452 (2005) A census of object types and redshift estimates in the SDSS photometric catalog from a trained decision tree classifier.

J/AJ/130/2453/  table1.dat table2.dat table3.dat  (20Kb)

Krisciunas K., Garnavich P.M., Challis P., Prieto J.L., Riess A.G., Barris B., Aguilera C., Becker A.C., Blondin S., Chornock R., Clocchiatti A., Covarrubias R., Filippenko A.V., Foley R.J., Hicken M., Jha S., Kirshner R.P., Leibundgut B., Li W., Matheson T., Miceli A., Miknaitis G., Rest A., Salvo M.E., Schmidt B.P., Smith R.C., Sollerman J., Spyromilio J., Stubbs C.W., Suntzeff N.B., Tonry J.L., Wood-Vasey W.M. Astron. J., 130, 2453-2472 (2005) Hubble space telescope observations of nine high-redshift ESSENCE supernovae.

J/AJ/130/2473/  table1.dat table2.dat  (133Kb)

Kovalev Y.Y., Kellermann K.I., Lister M.L., Homan D.C., Vermeulen R.C., Cohen M.H., Ros E., Kadler M., Lobanov A.P., Zensus J.A., Kardashev N.S., Gurvits L.I., Aller M.F., Aller H.D. Astron. J., 130, 2473-2505 (2005) Sub-milliarcsecond imaging of quasars and active galactic nuclei. IV. Fine-scale structure.

J/AJ/130/2529/  table3.dat table4.dat table5.dat  (18Kb)

Ojha R., Fey A.L., Charlot P., Jauncey D.L., Johnston K.J., Reynolds J.E., Tzioumis A.K., Quick J.F.H., Nicolson G.D., Ellingsen S.P., McCulloch P.M., Koyama Y. Astron. J., 130, 2529-2540 (2005) VLBI observations of southern hemisphere ICRF sources. II. Astrometric suitability based on intrinsic structure.

J/AJ/130/2541/  notes.dat table1.dat table2.dat table4.dat  (47Kb)

Miller N.A. Astron. J., 130, 2541-2558 (2005) Star formation and active galactic nuclei in the core of the Shapley supercluster: a Very Large Array survey of A3556, A3558, SC 1327-312, SC 1329-313, and A3562.

J/AJ/130/2571/  table2.dat table4.dat  (39Kb)

Jangren A., Salzer J.J., Sarajedini V.L., Gronwall C., Werk J.K., Chomiuk L., Moody J.W., Boroson T.A. Astron. J., 130, 2571-2583 (2005) The KPNO international spectroscopic survey. V. Hα-selected survey list 3.

J/AJ/130/2584/  table1.dat table2.dat table3.dat table4.dat table5.dat  (86Kb)

Salzer J.J., Jangren A., Gronwall C., Werk J.K., Chomiuk L.B., Caperton K.A., Melbourne J., Mckinstry K. Astron. J., 130, 2584-2597 (2005) Spectroscopy of KISS emission-line galaxy candidates. IV. KPNO and WIYN observations.

J/AJ/130/2613/  notes.dat table2.dat table3.dat table4.dat  (53Kb)

Giovanelli R., Haynes M.P., Kent B.R., Perillat P., Catinella B., Hoffman G.L., Momjian E., Rosenberg J.L., Saintonge A., Spekkens K., Stierwalt S., Brosch N., Masters K.L., Springob C.M., Karachentsev I.D., Karachentseva V.E., Koopmann R.A., Muller E., van Driel W., van Zee L. Astron. J., 130, 2613-2624 (2005) The Arecibo Legacy Fast ALFA survey. II. Results of precursor observations.

J/AJ/130/2647/  table1.dat  (12Kb)

van Dokkum P.G. Astron. J., 130, 2647-2665 (2005) The recent and continuing assembly of field elliptical galaxies by red mergers.

J/AJ/130/2666/  table2.dat table3.dat  (28Kb)

Prochaska L.C., Rose J.A., Schiavon R.P. Astron. J., 130, 2666-2676 (2005) A new definition for the Ca4227 feature: is calcium really underabundant in early-type galaxies?

J/AJ/130/2677/  table3.dat table5.dat  (22Kb)

Majewski S.R., Frinchaboy P.M., Kunkel W.E., Link R., Munoz R.R., Ostheimer J.C., Palma C., Patterson R.J., Geisler D. Astron. J., 130, 2677-2700 (2005) Exploring halo substructure with giant stars. VI. Extended distributions of giant stars around the Carina dwarf spheroidal galaxy: how reliable are they?

J/AJ/130/2701/  plates/* table3.dat  (45.6 Mb)

Fresneau A., Vaughan A.E., Argyle R.W. Astron. J., 130, 2701-2716 (2005) Density wave streaming motions in stars along the Sagittarius spiral arm.

J/AJ/130/2804/  refs.dat table1.dat table3.dat table4.dat table6.dat table7.dat  (55Kb)

Rossi S., Beers T.C., Sneden C., Sevastyanenko T., Rhee J., Marsteller B. Astron. J., 130, 2804-2823 (2005) Estimation of carbon abundances in metal-poor stars. I. Application to the strong G-band stars of Beers, Preston, and Shectman.

J/AJ/130/2838/  table1.dat table2.dat table4.dat  (86Kb)

Lacy C.H.S., Torres G., Claret A., Vaz L.P.R. Astron. J., 130, 2838-2846 (2005) Absolute properties of the eclipsing binary star RW Lacertae.

J/AJ/130/2866/  table1.dat table2.dat table3.dat table4.dat table5.dat table6.dat table7.dat  (33Kb)

Muterspaugh M.W., Lane B.F., Konacki M., Burke B.F., Colavita M.M., Kulkarni S.R., Shao M. Astron. J., 130, 2866-2875 (2005) PHASES high-precision differential astrometry of δEquulei.

J/AJ/131/1/  table2.dat table3.dat table4.dat table8.dat table9.dat table10.dat  (165Kb)

Hennawi J.F., Strauss M.A., Oguri M., Inada N., Richards G.T., Pindor B., Schneider D.P., Becker R.H., Gregg M.D., Hall P.B., Johnston D.E., Fan X., Burles S., Schlegel D.J., Gunn J.E., Lupton R.H., Bahcall N.A., Brunner R.J., Brinkmann J. Astron. J., 131, 1-23 (2006) Binary quasars in the Sloan Digital Sky Survey: evidence for excess clustering on small scales.

J/AJ/131/49/  table1.dat table2.dat  (22Kb)

Richards G.T., Haiman Z., Pindor B., Strauss M.A., Fan X., Eisenstein D., Schneider D.P., Bahcall N.A., Brinkmann J., Fukugita M. Astron. J., 131, 49-54 (2006) A snapshot survey for gravitational lenses among z =4.0 quasars. II. Constraints on the 4.0 z 5.4 quasar population.

J/AJ/131/84/  table2.dat  (3.3 Mb)

Vanden Berk D.E., Shen J., Yip C.-W., Schneider D.P., Connolly A.J., Burton R.E., Jester S., Hall P.B., Szalay A.S., Brinkmann J. Astron. J., 131, 84-99 (2006) Spectral decomposition of broad-line AGNs and host galaxies.

J/AJ/131/100/  refs4.dat refs5.dat table4.dat table5.dat  (65Kb)

Burgess A.M., Hunstead R.W. Astron. J., 131, 100-113 (2006) The Molonglo Southern 4Jy sample (MS4). I. Definition.

J/AJ/131/114/  refs.dat table1.dat table2.dat table3.dat table7.dat  (64Kb)

Burgess A.M., Hunstead R.W. Astron. J., 131, 114-132 (2006) The Molonglo Southern 4Jy sample (MS4). II. ATCA imaging and optical identification.

J/AJ/131/143/ table1.dat table2.dat table3.dat table4.dat table5.dat table6.dat table7.dat table8.dat (33Kb) Homeier N.L., Postman M., Menanteau F., Blakeslee J.P., Mei S., Demarco R., Ford H.C., Illingworth G.D., Zirm A. Astron. J., 131, 143-157 (2006) Imprints of environment on cluster and field late-type galaxies at z 1.

J/AJ/131/250/  table2.dat table3.dat  (62Kb)

Frayer D.T., Fadda D., Yan L., Marleau F.R., Choi P.I., Helou G., Soifer B.T., Appleton P.N., Armus L., Beck R., Dole H., Engelbracht C.W., Fang F., Gordon K.D., Heinrichsen I., Henderson D., Hesselroth T., Im M., Kelly D.M., Lacy M., Laine S., Latter W.B., Mahoney W., Makovoz D., Masci F.J., Morrison J.E., Moshir M., Noriega-Crespo A., Padgett D.L., Pesenson M., Shupe D.L., Squires G.K., Storrie-Lombardi L.J., Surace J.A., Teplitz H.I., Wilson G. Astron. J., 131, 250-260 (2006) Spitzer 70 and 160μm observations of the extragalactic First Look Survey.

J/AJ/131/332/  table3.dat  (52Kb)

Geha M., Guhathakurta P., Rich R.M., Cooper M.C. Astron. J., 131, 332-342 (2006) Local Group dwarf elliptical galaxies. I. Mapping the dynamics of NGC 205 beyond the tidal radius.

J/AJ/131/407/  table1.dat table2.dat  (292Kb)

Kaluzny J., Mochnacki S., Rucinski S.M. Astron. J., 131, 407-413 (2006) Variable stars in the Large Magellanic Cloud: discovery of extragalactic W UMa binaries.

J/AJ/131/414/  table1.dat table2.dat  (46Kb)

Hill A., Zaritsky D. Astron. J., 131, 414-430 (2006) The star clusters of the Small Magellanic Cloud: structural parameters.

J/AJ/131/431/ table1.dat table2.dat table3.dat table5.dat table6.dat table10.dat table11.dat (46Kb) Brewer M.-M, Carney B.W. Astron. J. 131, 431 (2006) A comparison of the chemical evolutionary histories of the galactic thin disk and thick disk stellar populations.

J/AJ/131/461/  table1.dat  (131Kb)

Anthony-Twarog B.J., Tanner D., Cracraft M., Twarog B.A. Astron. J., 131, 461-472 (2006) vbyCaHβCCD photometry of clusters. VI. The metal-deficient open cluster NGC 2420.

J/AJ/131/527/  sn.dat table2.dat table4.dat table5.dat  (110Kb)

Jha S., Kirshner R.P., Challis P., Garnavich P.M., Matheson T., Soderberg A.M., Graves G.J.M., Hicken M., Alves J.F., Arce H.G., Balog Z., Barmby P., Barton E.J., Berlind P., Bragg A.E., Briceno C., Brown W.R., Buckley J.H., Caldwell N., Calkins M.L., Carter B.J., Concannon K.D., Donnelly R.H., Eriksen K.A., Fabricant D.G., Falco E.E., Fiore F., Garcia M.R., Gomez M., Grogin N.A., Groner T., Groot P.J., Haisch K.E.Jr, Hartmann L., Hergenrother C.W., Holman M.J., Huchra J.P., Jayawardhana R., Jerius D., Kannappan S.J., Kim D.-W., Kleyna J.T., Kochanek C.S., Koranyi D.M., Krockenberger M., Lada C.J., Luhman K.L., Luu J.X., Macri L.M., Mader J.A., Mahdavi A., Marengo M., Marsden B.G., McLeod B.A., McNamara B.R., Megeath S.T., Moraru D., Mossman A.E., Muench A.A., Munoz J.A., Muzerolle J., Naranjo O., Nelson-Patel K., Pahre M.A., Patten B.M., Peters J., Peters W., Raymond J.C., Rines K., Schild R.E., Sobczak G.J., Spahr T.B., Stauffer J.R., Stefanik R.P., Szentgyorgyi A.H., Tollestrup E.V., Vaisanen P., Vikhlinin A., Wang Z., Willner S.P., Wolk S.J., Zajac J.M., Zhao P., Stanek K.Z. Astron. J. 131, 527 (2006) UBVRI light curves of 44 type Ia supernovae

J/AJ/131/582/  table1.dat table2.dat  (24Kb)

Kilic M., Munn J.A., Harris H.C., Liebert J., von Hippel T., Williams K.A., Metcalfe T.S., Winget D.E., Levine S.E. Astron. J., 131, 582-599 (2006) Cool white dwarfs in the Sloan Digital Sky Survey.

J/AJ/131/612/  table1.dat  (14Kb)

Smith B.J., Price S.D., Moffett A.J. Astron. J., 131, 612-620 (2006) Phase lags in the optical-infrared light curves of asymptotic giant branch stars.

J/AJ/131/621/  lc/* tab1.dat tab1.fit  (10.5 Mb)

Gettel S.J., Geske M.T., McKay T.A. Astron. J., 131, 621-632 (2006) A catalog of 1022 bright contact binary stars.

J/AJ/131/633/  table1.dat  (15Kb)

Geske M.T., Gettel S.J., McKay T.A. Astron. J., 131, 633-637 (2006) A ROSAT survey of contact binary stars.

J/AJ/131/666/  table4.dat  (14Kb)

de Vries W.H., Becker R.H., White R.L. Astron. J., 131, 666-679 (2006) Double-lobed radio quasars from the Sloan Digital Sky Survey.

J/AJ/131/716/  table2.dat  (18Kb)

Koopmann R.A., Haynes M.P., Catinella B. Astron. J. 131, 716 (2006) A comparison of Hαand stellar scale lengths in Virgo and field spirals.

J/AJ/131/814/  table1.dat table3.dat  (12Kb)

Beasley M.A., Strader J., Brodie J.P., Cenarro A.J., Geha M. Astron. J., 131, 814-827 (2006) The globular cluster system of the Virgo dwarf elliptical galaxy VCC 1087.

J/AJ/131/866/  table2.dat table4.dat table5.dat table6.dat  (29Kb)

Sohn S.T., O’Connell R.W., Kundu A., Landsman W.B., Burstein D., Bohlin R.C., Frogel J.A., Rose J.A. Astron. J., 131, 866-888 (2006) Hot populations in M87 globular clusters.

J/AJ/131/895/  table6.dat  (97Kb)

Koch A., Grebel E.K., Wyse R.F.G., Kleyna J.T., Wilkinson M.I., Harbeck D.R., Gilmore G.F., Evans N.W. Astron. J., 131, 895-911 (2006) Complexity on small scales: the metallicity distribution of the Carina dwarf spheroidal galaxy.

J/AJ/131/912/  table2.dat table3.dat  (907Kb)

Olszewski E.W., Mateo M., Harris J., Walker M.G., Coleman M.G., Da Costa G.S. Astron. J. 131, 912 (2006) The nature of the density clump in the Fornax dwarf spheroidal galaxy.

J/AJ/131/951/  table2.dat  (10Kb)

Roman-Lopes A., Abraham Z. Astron. J., 131, 951-959 (2006) The young massive stellar cluster associated with RCW 121.

J/AJ/131/1000/  objects.dat table3.dat  (63Kb)

Horch E.P., Davidson J.W.JR, Van Altena W.F., Girard T.M., Lopez C.E., Franz O.G., Timothy J.G. Astron. J., 131, 1000-1006 (2006) Speckle interferometry of southern double stars. III. Measures from the Cesco observatory, 1994-1996.

J/AJ/131/1015/  table5.dat  (9Kb)

Gatewood G., Han I. Astron. J., 131, 1015-1021 (2006) An Astrometric Study of Procyon.

J/AJ/131/1032/  table2.dat table3.dat table4.dat table6.dat  (24Kb)

Qian S.-B., Zhu L.-Y. Astron. J., 131, 1032-1043 (2006) BO Canum Venaticorum and SS Comae Berenices: a photometric study of AW UMa-type binaries.

J/AJ/131/1044/  table1.dat  (219Kb)

Drake A.J. Astron. J., 131, 1044-1056 (2006) Chromospherically active stars in the Galactic Bulge.

J/AJ/131/1184/  table1.dat table2.dat table3.dat  (11Kb)

Maiz-Apellaniz J. Astron. J., 131, 1184-1199 (2006) A recalibration of optical photometry: Tycho-2, Stromgren, and Johnson systems.

J/AJ/131/1216/  table1.dat table2.dat  (25Kb)

Afonso J., Mobasher B., Koekemoer A., Norris R.P., Cram L. Astron. J., 131, 1216-1230 (2006) Optical and X-ray identification of faint radio sources in the GOODS CDF-S Advanced Camera for Surveys field.

J/AJ/131/1236/  table1.dat table3.dat table4.dat  (32Kb)

Dong X.Y., De Robertis M.M. Astron. J., 131, 1236-1252 (2006) Low-luminosity active galaxies and their central black holes.

J/AJ/131/1262/  table1.dat table2.dat table3.dat table5.dat  (23Kb)

Homan D.C., Lister M.L.

Astron. J., 131, 1262-1279 (2006) MOJAVE: Monitoring of Jets in Active Galactic Nuclei with VLBA Experiments. II. First-epoch 15GHz circular polarization results.

J/AJ/131/1280/  table1.dat table2.dat table3.dat  (24Kb)

Fleenor M.C., Rose J.A., Christiansen W.A., Johnston-Hollitt M., Hunstead R.W., Drinkwater M.J., Saunders W. Astron. J., 131, 1280-1287 (2006) Redshifts and velocity dispersions of galaxy clusters in the Horologium-Reticulum supercluster.

J/AJ/131/1288/  sp/* table1.dat table2.dat table3.dat table4.dat table7.dat table8.dat  (48.3 Mb)

Bernardi M., Nichol R.C., Sheth R.K., Miller C.J., Brinkmann J. Astron. J., 131, 1288-1317 (2006) Evolution and environment of early-type galaxies.

J/AJ/131/1416/  table1.dat table2.dat table3.dat  (47Kb)

Nantais J.B., Huchra J.P., Barmby P., Olsen K.A.G., Jarrett T.H. Astron. J., 131, 1416-1425 (2006) Nearby spiral globular cluster systems. I. Luminosity functions.

J/AJ/131/1501/  table1.dat table2.dat  (20Kb)

Kerton C.R., Knee L.B.G., Schaeffer A.J. Astron. J., 131, 1501-1514 (2006) Intermediate-velocity gas in the Canadian Galactic Plane Survey.

J/AJ/131/1530/  table3.dat  (10Kb)

Herbig G.H., Dahm S.E. Astron. J., 131, 1530-1543 (2006) The pre-main-sequence population of L988.

J/AJ/131/1559/  table1.dat  (33Kb)

van den Bergh S. Astron. J., 131, 1559-1564 (2006) Diameters of open star clusters.

J/AJ/131/1565/  table2.dat table3.dat  (90Kb)

Krusberg Z.A.C., Chaboyer B. Astron. J., 131, 1565-1573 (2006) UBVI CCD photometry of the old open cluster Berkeley 17.

J/AJ/131/1574/  table1.dat table2.dat  (50Kb)

Lada C.J., Muench A.A., Luhman K.L., Allen L., Hartmann L., Megeath T., Myers P., Fazio G., Wood K., Muzerolle J., Rieke G., Siegler N., Young E. Astron. J., 131, 1574-1607 (2006) Spitzer observations of IC 348: the disk population at 2-3 million years.

J/AJ/131/1608/  table3.dat  (122Kb)

Oasa Y., Tamura M., Nakajima Y., Itoh Y., Maihara T., Iwamuro F., Motohara K., Hayashi S.S., Hayashi M., Kaifu N. Astron. J., 131, 1608-1628 (2006) Very low luminosity young cluster and the luminosity and mass functions in S106.

J/AJ/131/1648/  refs.dat table2.dat table3.dat table5.dat table6.dat table7.dat table8.dat  (58Kb)

Blondin S., Dessart L., Leibundgut B., Branch D., Hoflich P., Tonry J.L., Matheson T., Foley R.J., Chornock R., Filippenko A.V., Sollerman J., Spyromilio J., Kirshner R.P., Wood-Vasey W.M., Clocchiatti A., Aguilera C., Barris B., Becker A.C., Challis P., Covarrubias R., Davis T.M., Garnavich P., Hicken M., Jha S., Krisciunas K., Li W., Miceli A., Miknaitis G., Pignata G., Prieto J.L., Rest A., Riess A.G., Salvo M.E., Schmidt B.P., Smith R.C., Stubbs C.W., Suntzeff N.B. Astron. J., 131, 1648-1666 (2006) Using line profiles to test the fraternity of type Ia supernovae at high and low redshifts.

J/AJ/131/1674/  table1.dat table4.dat table5.dat  (287Kb)

Silvestri N.M., Hawley S.L., West A.A., Szkody P., Bochanski J.J., Eisenstein D.J., McGehee P., Schmidt G.D., Smith J.A., Wolfe M.A., Harris H.C., Kleinman S.J., Liebert J., Nitta A., Barentine J.C., Brewington H.J., Brinkmann J., Harvanek M., Krzesinski J., Long D., Neilsen E.H.Jr, Schneider D.P., Snedden S.A. Astron. J., 131, 1674-1686 (2006) A catalog of spectroscopically selected close binary systems from the Sloan Digital Sky Survey data release four.

J/AJ/131/1702/  table1.dat table3.dat  (17Kb)

Torres G.

Astron. J., 131, 1702-1711 (2006) The multiple system HD 27638.

J/AJ/131/1766/  table1.dat table2.dat table34.dat  (256Kb)

Carretta E. Astron. J., 131, 1766-1783 (2006) Abundances in red giant stars of NGC 2808 and correlations between chemical anomalies and global parameters in globular clusters.

J/AJ/131/1872/  table2.dat  (39Kb)

Petrov L., Kovalev Y.Y., Fomalont E.B., Gordon D. Astron. J., 131, 1872-1879 (2006) The fourth VLBA calibrator survey: VCS4.

J/AJ/131/1891/  catalog.dat zph.dat  (943Kb)

Foerster-Schreiber N.M., Franx M., Labbe I., Rudnick G., Van Dokkum P.G., Illingworth G.D., Kuijken K., Moorwood A.F.M., Rix H.-W., Roettgering H., Van Der Werf P. Astron. J., 131, 1891-1913 (2006) Faint infrared extragalactic survey: data and source catalog of the MS 1054-03 field.

J/AJ/131/1923/  table1.dat  (82Kb)

Rengstorf A.W., Brunner R.J., Wilhite B.C. Astron. J., 131, 1923-1933 (2006) A synoptic multiwavelength analysis of a large quasar sample.

J/AJ/131/1942/  table1.dat  (19Kb)

Chen P.S., Zhang P. Astron. J., 131, 1942-1947 (2006) Infrared photometric study for the IRAS 1Jy sample of ultraluminous infrared galaxies.

J/AJ/131/1948/  table2.dat  (8Kb)

Whalen D.J., Laurent-Muehleisen S.A., Moran E.C., Becker R.H. Astron. J., 131, 1948-1960 (2006) Optical properties of radio-selected narrow-line Seyfert 1 galaxies.

J/AJ/131/2018/  table1.dat table2.dat  (15Kb)

Bernardi M., Sheth R.K., Nichol R.C., Miller C.J., Schlegel D., Frieman J., Schneider D.P., Subbarao M., York D.G., Brinkmann J. Astron. J., 131, 2018-2034 (2006) A search for the most massive galaxies: double trouble?

J/AJ/131/2035/  table2.dat  (11Kb)

Galaz G., Villalobos A., Infante L., Donzelli C. Astron. J., 131, 2035-2049 (2006) Bulge evolution in face-on spiral and low surface brightness galaxies.

J/AJ/131/2089/  table5.dat table6.dat  (15Kb)

Sluis A.P.N., Williams T.B. Astron. J., 131, 2089-2104 (2006) Uncovering planetary nebulae in early-type galaxies using the Rutgers Fabry-Perot.

J/AJ/131/2114/  table1.dat table2.dat  (39Kb)

Walker M.G., Mateo M., Olszewski E.W., Bernstein R., Wang X., Woodroofe M. Astron. J., 131, 2114-2139 (2006) Internal kinematics of the Fornax dwarf spheroidal galaxy.

J/AJ/131/2164/  table1.dat table2.dat table3.dat table4.dat table5.dat  (52Kb)

Townsley L.K., Broos P.S., Feigelson E.D., Garmire G.P., Getman K.V. Astron. J., 131, 2164-2184 (2006) A Chandra ACIS study of 30 Doradus. II. X-ray point sources in the massive star cluster R136 and beyond.

J/AJ/131/2245/  table1.dat table2.dat  (21Kb)

Zhang T., Wang X., Li W., Zhou X., Ma J., Jiang Z., Chen J. Astron. J., 131, 2245-2255 (2006) Optical photometry of the type II-p supernova 2004dj in NGC 2403.

J/AJ/131/2274/  refs.dat table1.dat table4.dat  (30Kb)

Sozzetti A., Yong D., Carney B.W., Laird J.B., Latham D.W., Torres G. Astron. J., 131, 2274-2289 (2006) Chemical composition of the planet-harboring star TrES-1.

J/AJ/131/2373/  table4.dat table5.dat table6.dat  (116Kb)

Virani S.N., Treister E., Urry C.M., Gawiser E. Astron. J., 131, 2373-2382 (2006) The extended Chandra Deep Field-South survey: X-ray point-source catalog.

J/AJ/131/2383/  table1.dat table2.dat  (448Kb)

Glazebrook K., Verma A., Boyle B., Oliver S., Mann R.G., Monbleau D. Astron. J., 131, 2383-2393 (2006) Anglo-Australian Telescope imaging and microslit spectroscopy in the southern Hubble Deep Field.

J/AJ/131/2478/  refs.dat table4.dat table5.dat table8.dat table9.dat  (65.4 Mb)

Massey P., Olsen K.A.G., Hodge P.W., Strong S.B., Jacoby G.H., Schlingman W., Smith R.C. Astron. J., 131, 2478-2496 (2006) A survey of Local Group galaxies currently forming stars. I. UBVRI photometry of stars in M31 and M33.

J/AJ/131/2538/  table3.dat  (324Kb)

Lewis M.S., Liu W.M., Paust N.E.Q., Chaboyer B. Astron. J., 131, 2538-2542 (2006) A new color-magnitude diagram for Palomar 11.

J/AJ/131/2722/  table1.dat table2.dat table3.dat table4.dat table5.dat  (44Kb)

Chiu K., Fan X., Leggett S.K., Golimowski D.A., Zheng W., Geballe T.R., Schneider D.P., Brinkmann J. Astron. J., 131, 2722-2736 (2006) Seventy-one new L and T dwarfs from the Sloan Digital Sky Survey.

J/AJ/132/50/  table1.dat table2.dat table5.dat  (7.7 Mb)

Wycoff G.L., Mason B.D., Urban S.E. Astron. J., 132, 50-60 (2006) Data mining for double stars in astrometric catalogs.

J/AJ/132/85/  objects.dat table5.dat  (339Kb)

Preston G.W., Sneden C., Thompson I.B., Shectman S.A., Burley G.S. Astron. J., 132, 85-110 (2006) Atmospheres, chemical compositions, and evolutionary histories of very metal-poor red horizontal-branch stars in the galactic field and in NGC 7078 (M15).

J/AJ/132/111/  table2.dat  (12Kb)

Joner M.D., Taylor B.J., Laney C.D., Van Wyk F. Astron. J., 132, 111-116 (2006) Homogeneous photometry for the Hyades: scale-factor and zero-point tests of previously published BV(RI)_{C}photometry.

J/AJ/132/137/  table1.dat tables.dat  (91Kb)

Cohen J.G., McWilliam A., Shectman S., Thompson I., Christlieb N., Melendez J., Ramirez S., Swensson A., Zickgraf F.-J. Astron. J., 132, 137-160 (2006) Carbon stars in the Hamburg/ESO survey: abundances.

J/AJ/132/161/  notes.dat table2.dat table5.dat  (194Kb)

Gray R.O., Corbally C.J., Garrison R.F., McFadden M.T., Bubar E.J., McGahee C.E., O’Donoghue A.A., Knox E.R. Astron. J., 132, 161-170 (2006) Contributions to the nearby stars (NStars) project: spectroscopy of stars earlier than M0 within 40 pc. The southern sample.

J/AJ/132/197/  table1.dat table2.dat table3.dat  (41Kb)

Woods D.F., Geller M.J., Barton E.J. Astron. J., 132, 197-209 (2006) Tidally triggered star formation in close pairs of galaxies: major and minor interactions.

J/AJ/132/231/  fits/* sp/* table2.dat table3.dat table4.dat table5.dat  (364.2 Mb)

Papovich C., Cool R., Eisenstein D., Le Floc’h E., Fan X., Kennicutt R.C.JR, Smith J.D.T., Rieke G.H., Vestergaard M. Astron. J., 132, 231-241 (2006) An MMT hectospec redshift survey of 24 μm sources in the Spitzer first look survey.

J/AJ/132/242/  table2.dat table3.dat table4.dat table5.dat table6.dat  (24Kb)

Nord M.E., Henning P.A., Rand R.J., Lazio T.J.W., Kassim N.E. Astron. J., 132, 242-252 (2006) 74 MHz discrete H II absorption regions toward the inner galaxy.

J/AJ/132/271/  table3.dat  (45Kb)

Olsen K.A.G., Blum R.D., Stephens A.W., Davidge T.J., Massey P., Strom S.E., Rigaut F. Astron. J., 132, 271-289 (2006) The star formation histories of the bulge and disk of M31 from resolved stars in the near-infrared.

J/AJ/132/321/  table1.dat  (11Kb)

Deo R.P., Crenshaw D.M., Kraemer S.B. Astron. J., 132, 321-346 (2006) The host galaxies of narrow-line Seyfert 1 galaxies: nuclear dust morphology and starburst rings.

J/AJ/132/347/  table2.dat table3.dat table4.dat  (21Kb)

Chiboucas K., Mateo M. Astron. J., 132, 347-359 (2006) The luminosity function of nearby galaxy clusters. II. Redshifts and luminosity function for galaxies in the region of the Centaurus cluster.

J/AJ/132/378/  table3.dat table4.dat  (236Kb)

Bianchi L., Efremova B.V. Astron. J., 132, 378-392 (2006) The recent star formation in NGC 6822 from Hubble space telescope imaging.

J/AJ/132/393/  table2.dat table3.dat  (9Kb)

Hintz E.G., Rose M.B., Bush T.C., Maxwell A.A. Astron. J., 132, 393-400 (2006) Establishing observational baselines for two δScuti variables: V966 Herculis and V1438 Aquilae.

J/AJ/132/447/  clusters.dat photdata.dat smdata.dat  (149Kb)

Noyola E., Gebhardt K. Astron. J. 132, 447 (2006) Surface Brightness Profiles of Galactic Globular Clusters from Hubble Space Telescope Images.

J/AJ/132/1445/  appen.pdf spectra/* table1.dat table2.dat table3.dat  (2.7 Mb)

Engelke C.W., Price S.D., Kraemer K.E. Astron. J., 132, 1445-1463 (2006) Spectral irradiance calibration in the infrared. XVI. Improved accuracy in the infrared spectra of the secondary and tertiary standard calibration stars.

J/AJ/133/439/ refs.dat table2.dat table3.dat table4.dat table5.dat table6.dat table9.dat table10.dat (65Kb) Cruz K.L., Reid I.N., Kirkpatrick J.D., Burgasser A.J., Liebert J., Solomon A.R., Schmidt S.J., Allen P.R., Hawley S.L., Covey K.R. Astron. J., 133, 439-467 (2007) Meeting the cool neighbors. IX. The luminosity function of M7-L8 ultracool dwarfs in the field.

J/AJ/133/1470/  lc/* plottsvsc.m tsvsc1.dat  (4.3 Mb)

Damerdji Y., Klotz A., Boer M. Astron. J., 133, 1470-1477 (2007) The TAROT suspected variable star catalog.

J/AJ/133/1947/  table1.dat table2.dat  (193Kb)

Nieppola E., Tornikoski M., Lahteenmaki A., Valtaoja E., Hakala T., Hovatta T., Kotiranta M., Nummila S., Ojala T., Parviainen M., Ranta M., Saloranta P.-M., Torniainen I., Troller M. Astron. J., 133, 1947-1953 (2007) 37 GHz observations of a large sample of BL Lacertae objects.

J/AJ/133/2258/  refs.dat table1.dat table2.dat  (47Kb)

Schmidt S.J., Cruz K.L., Bongiorno B.J., Liebert J., Reid I.N. Astron. J., 133, 2258-2273 (2007) Activity and kinematics of ultracool dwarfs, including an amazing flare observation.

J/AJ/134/1103/  cw1255.dat ehdfs1.dat ehdfs2.dat sdss1030.dat  (4.4 Mb)

Quadri R., Marchesini D., Van Dokkum P., Gawiser E., Franx M., Lira P., Rudnick G., Urry C.M., Maza J., Kriek M., Barrientos L.F., Blanc G.A., Castander F.J., Christlein D., Coppi P.S., Hall P.B., Herrera D., Infante L., Taylor E.N., Treister E., Willis J.P. Astron. J., 134, 1103-1117 (2007) The Multiwavelength Survey by Yale-Chile (MUSYC): deep near-infrared imaging and the selection of distant galaxies.

J/AJ/134/1245 See VIII/79/

J/AJ/162/38 See II/267/

J/AN/283/109/  table.dat  (77Kb)

Heckmann O., Dieckvoss W., Kox H. Astron. Nachrichten, 283, 109-139 (1956) Eigenbewegungen in der Umgebung von alpha Persei.

J/AN/316/225/  table6  (98Kb)

Andruk V., Kharchenko N., Schilbach E., Scholz R.-D. Astron. Nachrichten, 316, 225 (1995) Photometric survey near the main Galactic meridian: 1. Photoelectric stellar magnitudes and colours in the UBVR system

J/AN/318/35/  newpn.dat notes1.dat notes2.dat paper.ps postpn.dat prepn.dat refs.dat  (183Kb)

Kohoutek L. Astron. Nachrichten, 318, 35 (1997) New and misclassified planetary nebulae

J/AN/319/201/  notes.dat table1.dat table2.dat  (28Kb)

Prosser C.F., Randich S Astron. Nachrichten 319, 201 (1998) Identification of new low-mass members of the alpha Persei open cluster by ROSAT. I.

J/AN/319/215/  appena.dat table1.dat table2.dat table4.dat table7.dat  (46Kb)

Prosser C.F., Randich S., Simon T. Astron. Nachrichten, 319, 215 (1998) Identification of new low-mass members of the alpha Persei open cluster by ROSAT. II.

J/AN/321/81/  notes.dat table1.dat  (39Kb)

Lohmann P., Wendker H.J. Astron. Nachrichten, 321, 81 (2000) ROSAT PSPC observations of the Orion Trapezium area. II. Source variability and revised source list.

J/AN/321/101/  atlas/* ms.ps table2.dat table3.dat  (6.2 Mb)

Allam S.S., Tucker D.L. Astron. Nachrichten, 321, 101-114 (2000) Compact groups of galaxies in the Las Campanas Redshift Survey.

J/AN/321/277/  table5.dat table6.dat  (53Kb)

Strassmeier K.G., Schordan P. Astron. Nachrichten, 321, 277 (2000) A temperature calibration for MK-class III giants from high-resolution spectral line-depth ratios

J/AN/323/31/  table4.dat table5.dat  (148Kb)

Ekmekci F., Oezeren F.F., Ak H. Astron. Nachrichten, 323, 31 (2002) Chromospherically active binary systems, RT And and ER Vul: 1995-1998 observations and spot distributions.

J/AN/323/139/  notes.dat table1.dat table2.dat  (33Kb)

Sperauskas J., Bartkevicius A. Astron. Nachrichten, 323, 139 (2002) Radial velocities of population II binary stars. I.

J/AN/323/484/  remarks.dat table1.dat  (61Kb)

Kohoutek L., Kuehl D. Astron. Nachrichten, 323, 484 (2002) Accurate coordinates of Planetary Nebulae.

J/AN/324/219/  refs.dat table1.dat  (58Kb)

Wegner W. Astron. Nachrichten, 324, 219 (2003) The total-to-selective extinction ratio determined from near IR photometry of OB stars

J/AN/324/437/  notes.dat table1.dat table1a.dat table2.dat table3.dat table3a.dat  (24Kb)

Kohoutek L., Wehmeyer R. Astron. Nachr., 324, 437-444 (2003) New H-alpha emission stars towards the Galactic Bulge.

J/AN/325/307/  table2.dat  (16Kb)

Notni P., Karachentsev I.D., Makarova L.N. Astron. Nachr., 325, 307-316 (2004) The central region of NGC 3077: ground-based photometry and age of bright ‘starlike’ objects.

J/AN/325/424/  bmag.dat vmag.dat  (38Kb)

Patkos L., Csizmadia Sz. Astron. Nachrichten, 325, 424 (2004) The eclipsing active star SV Camelopardalis.

J/AN/325/705/  table2.dat  (15Kb)

Bejar V.J.S., Zapatero Osorio M.R., Rebolo R. Astron. Nachrichten, 325, 705-713 (2004) Optical and infrared photometry of new very low-mass stars and brown dwarfs in the sigma Orionis cluster.

J/AN/325/740/  cluster.dat csoca.dat  (38.2 Mb)

Kharchenko N.V., Piskunov A.E., Roeser S., Schilbach E., Scholz R.-D. Astron. Nachr., 325, 740-748 (2004) Astrophysical supplements to the ASCC-2.5. II. Membership probabilities in 520 Galactic open cluster sky areas.

J/AN/326/3/  table8.dat table9.dat table10.dat table11.dat  (269Kb)

Ahumada J.A. Astron. Nachr., 326, 3-18 (2005) CCD photometry of the open clusters NGC 2627, NGC 5617, Pismis 7, and Ruprecht 75.

J/AN/326/321/  table1.dat  (46Kb)

Bilir S., Karaali S., Tuncel S. Astronomische Nachrichten, 326, 321-331 (2005) Absolute magnitudes for late-type dwarf stars for Sloan photometry.

J/AN/327/365/  2dfgaln.dat 2dfgalni.dat 2dfgals.dat 2dfgalsi.dat 2dfgrn.dat 2dfgrs.dat  (23.2 Mb)

Tago E., Einasto J., Saar E., Einasto M., Suhhonenko I., Joeveer M., Vennik J., Heinamaki P., Tucker D.L. Astron. Nachr., 327, 365-378 (2006) Clusters and groups of galaxies in the 2dF galaxy redshift survey: a new catalogue.

J/AN/328/889/  crvoca.dat refs.dat  (54Kb)

Kharchenko N.V., Scholz R.-D., Piskunov A.E., Roeser S., Schilbach E. Astron. Nachrichten, 328, 889 (2007) Astrophysical supplements to the ASCC-2.5. Ia. Radial velocities of 55000 stars and mean radial velocities of 516 Galactic open clusters and associations.

J/AN/328/938/  table5.dat  (43Kb)

Biazzo K., Frasca A., Catalano S., Marilli E. Astron. Nachrichten, 328, 938-947 (2007) Effective temperature vs. line-depth ratio for ELODIE spectra: gravity and rotational velocity effects.

J/AZh/62/229/  table.dat  (8Kb)

Berlin A.B., Gol’nev V.Ya., Lipovka N.M., Nizhel’skij N.A., Spangenberg E.E. Astron. Zh., 62, 229 (1985) Results of observation made at Ratan-600 with the high-sensitivity radiometer at the wavelength 7.6cm of the region of the Galaxy in the interval alpha=17h30m-19h30m, delta=-25.21 -22.20 (1950)

J/AZh/71/37/  table1  (10Kb)

Slysh V.I., Bachiller R., Berulis I.I., Valtts I.E., Gomez-Gonzalez J., Kalenskii S.V. Astron. Zh. 71, 37 (1994) Detection of Thermal Methanol Emission from 33 Hot and Cold Clouds at 48 GHz

J/AZh/71/110/  table2  (25Kb)

Kovalenko A.V., Pynzar A.V., Udaltsov V.A. Astron. Zh. 71, 110 (1994) Observations of Supernova Remnants at Pushchino: Catalog of Flux Densities at Meter Wavelengths.

J/AZh/71/189/  table2.dat table3.dat  (10Kb)

Sergeev S.G. Astron. Zh. 71, 189 (1994) Gas kinematics: estimation of the broad-line region size and the central source mass from the profile variability of the Hαline in NGC 4151

J/AZh/71/200/  table  (15Kb)

Doroshenko V.T., Sergeev S. G., Chuvaev K. K. Astron. Zh. 71, 200 (1994) Spectral Investigation of NGC 7469 Using Image-Tube Observations from 1972-1990.

J/AZh/71/352/  table6  (9Kb)

Verkhodanov O.V.

Astron. Zh. 71, 352 (1994) Study of the Objects of the 1988 Zenith Sky Survey.

J/AZh/71/468/  table3  (12Kb)

Tereshchenko V.M. Astron. Zh. 71, 468 (1994) Regional Spectrophotometric Standards of Intermediate Brightness: Standards around the Northern Celestial Pole

J/AZh/71/572/  table2  (17Kb)

Shevchenko V. S. Astron. Zh. 71, 572 (1994) V380 Ori: An Ae/Be Herbig Star at an Early Stage of Its Envelope Formation. Emission Lines of Metals.

J/AZh/71/588/  table2  (14Kb)

Zakharova L.A. Astron. Zh. 71, 588 (1994) Study of Two HgMn Stars with Suspected Anomalies of Noble Gas Abundances.

J/AZh/71/618/  table  (5Kb)

Kolotilov E.A., Nadzhip A.E., Shenavrin V.I., Yudin B.F. Astron. Zh. 71, 618 (1994) UBVJHKLM Photometry of Nova Cygni 1992.

J/AZh/71/684/  table1 table2  (32Kb)

Soboleva N.S., Parijskij Yu.N., Naugolnaya M.N. Astron. Zh. 71, 684 (1994) Optical Identification of the Objects of RC and UTRAO Catalogs on Enlarged Palomar Sky Atlas Prints

J/AZh/71/706/  table1 table2 table3 table4 table5  (33Kb)

Silchenko O.K. Astron. Zh. 71, 706 (1994) Photometrically Distinct Nuclei in Elliptical and Early-Type Disk Galaxies.

J/AZh/71/717/  table  (17Kb)

Gagen-Torn V.A., Marchenko S.G., Yurchenko A.V., Yakovleva V.A. Astron. Zh. 71, 717 (1994) Polarimetric Observations of BL Lac

J/AZh/71/762/  table.dat table.ps table.tex  (32Kb)

Malov I.F., Malofeev V.M., Sen’e D.S. Astron. Zh. 71, 762 (1994) The observed integral radio luminosities of 232 pulsars.

J/AZh/72/12/  table  (18Kb)

Mingaliev M., Khabrakhmanov A. Astron. Zh. 72, 12 (1995) Spectroscopic Study of a Sample of Radio Sources from the 87GB Catalog

J/AZh/72/22/  table1  (7Kb)

Valtts I.E., Dzyura A.M., Kalenskii S.V., Slysh V.I., Bus R., Vinnberg A. Astron. Zh. 72, 22 (1995) A 95 GHz Survey of Methanol Masers

J/AZh/72/80/  table2  (8Kb)

Mironov A.V., Moshkalev V.G. Astron. Zh. 72, 80 (1995) Analysis of Spectrophotometric and Photometric Data for Primary Spectrophotometric Standards

J/AZh/72/291/  table1.dat  (18Kb)

Gorshkov, A.G., Konnikova, V. K. Astron. Zh. 72, 291 (1995) Variability at Frequencies 3.9 and 7.5 GHz: Radio Sources from the Zelenchukskii Survey with Fluxes 200mJy.

J/AZh/72/447/  table.dat  (14Kb)

Babadzhanyants M.K., Belokonj E.T., Gamm N.G. Astron. Zh. 72, 447 (1995) Optical Monitoring of the Superluminal Quasar 3C345 from 1984 to 1991

J/AZh/72/495/  table4.dat table5.dat table6.dat table7.dat table8.dat table9.dat  (24Kb)

Orlov V.V., Panchenko I.E., Rastorguev A.S., Yatsevich A.V. Astron. Zh. 72, 495 (1995)

Moving Star Clusters in the Solar Neighborhood

J/AZh/72/613/  notes.dat table2.dat  (19Kb)

Kopylov A.I., Goss V.M., Parijskij Yu.N., Soboleva N.S., Zhelenkova O.P., Tempirova A.V., Vitkovskii V.V., Naugolnaya M.N., Verkhodanov O. V. Astron. Zh. 72, 613 (1995) Optical identification of a subsample of steep spectrum radio sources from the RC Catalog using the 6-m SAO telescope: optical observations

J/AZh/72/855/  table2.dat  (9Kb)

Antipova L.I., Berdyugina S.V., Savanov I.S. Astron. Zh. 72, 855 (1995) On the chemical composition of the G7 IIIa red giant eta Piscium.

J/AZh/72/864/  table3.dat table3b.dat  (9Kb)

Boyarchuk A.A., Antipova L.I., Boyarchuk M.E., Savanov I.S. Astron. Zh. 72, 864 (1995) Elemental abundances in the atmosphere of the red giant gamma Sagittae.

J/AZh/73/67/  table2.dat  (17Kb)

Knyazeva L.N., Kharitonov A.V. Astron. Zh. 73, 67 (1996) On the normal spectral energy distribution of stars: G, K, and M stars.

J/AZh/73/194/  table  (10Kb)

Grinin V.P., Rostopchina A.N. Astron. Zh. 73, 194 (1996) Orientation of Circumstellar Disks and the Statistics of HαProfiles of Ae/Be Herbig Stars.

J/AZh/73/762/  table1.dat table2.dat table3.dat  (26Kb)

Tereshchenko V.M. Astron. Zh. 73, 762 (1996) The energy distribution in the visible spectrum for the 27 giants and supergiants of the M spectral type

J/AZh/73/835/  table.dat  (17Kb)

Fletcher A., Conner S., Crawford F., Cartwright J., Burke B., Pariiskii Yu.N., Soboleva N.S., Kopylov A.I.,Temirova A.V., Verkhodanov O.V., Naugol’naya M.N. Astron. Zh. 73, 835 (1996) Identification of the Objects from the RC catalog with accuracy exceeding 1” using the Palomar Prints and VLA maps from the MIT archive.

J/AZh/73/862/  table1.dat  (19Kb)

Boyarchuk A.A., Antipova L.I., Boyarchuk M.E., Savanov I.S. Astron. Zh. 73, 862 (1996) The atmospheres of six red giants

J/AZh/73/906/  table4.dat  (12Kb)

Knyazeva L.N., Kharitonov A.V. Astron. Zh. 73, 906 (1996) On the normal energy distribution in stellar spectra: main-sequence B stars.

J/AZh/74/42/  table3.dat  (44Kb)

Bursov N.N Astron. Zh., 74, 42 (1997) Spectra of weak radio sources from the RATAN-600 ”Kholod” deep survey

J/AZh/74/63/  table1.dat  (29Kb)

Malov I.F., Malov O.I. Astron. Zh. 74, 63 (1997) Cluster analysis of pulsar parameters

J/AZh/74/812/  table1.dat  (13Kb)

Tyul’bashev S.A. Astron. Zh., 74, 812 (1997) Analysis of compact radio sources using 102 MHz interplanetary scintillation observations

J/AZh/74/853/  refs.dat table1.dat table2.dat  (46Kb)

Mal’kov Yu.F. Astron. Zh. 74, 853 (1997) A self-consistent determination of the distances, physical parameters and chemical composition for a large sample of galactic planetary nebulae: the distances and parameters of central stars and the optical depth of

envelopes.

J/AZh/75/70/  table2.dat table3.dat table6.dat  (30Kb)

Knyazeva L.N., Kharitonov A.V. Astron. Zh. 75, 70 (1998) On the normal spectral energy distribution of stars: Spectral types O9-B5.

J/AZh/75/197/  table1.dat table4.dat  (9Kb)

Knyazeva L.N., Kharitonov A.V. Astron. Zh. 75, 197 (1998) On the normal spectral energy distribution in stellar spectra: O5-O8 stars.

J/AZh/75/231/  table.dat  (8Kb)

Pavlenko E.P. Astron. Zh. 75, 231 (1998) MV Lyr in low brightness states. 2. Bursts and periods

J/AZh/75/586/  table1.dat table1b.dat  (19Kb)

Boyarchuk A.A., Antipova L.I., Boyarchuk M.E., Savanov I.S. Astron. Zh. 75, 586 (1998) Studies of the Sun as a standard for analysis of stellar atmospheric abundances

J/AZh/76/83/  refs.dat table1.dat table2.dat  (31Kb)

Botashev A.M., Verkhodanov O.V., Cardona O., Lipovka N.M., Lipovka A.A., Mingaliev M.G., Chavira E. Astron. Zh. 76, 83 (1999) Investigation of galaxies in and near the cluster A1185.

J/AZh/76/604/  stars.dat table3.dat  (17Kb)

Tereshchenko V.M. Astron. Zh. 76, 604 (1999) Spectrophotometry of three S stars and thirteen carbon N stars.

J/AZh/77/3/  table.dat  (16Kb)

Lipovka N.M., Lipovka A.A., Verkhodanov O.V., Chavira E. Astron. Zh. 77, 3 (2000) Large studies of bright steep-Spectrum radio sources.

J/AZh/77/21/  table2.dat table3.dat  (65Kb)

Dagkesamanskii R.D., Samodurov V.A., Lapaev K.A. Astron. Zh. 77, 21 (2000) A sky survey at 102.5MHz: radio sources at declinations 27.5^{∘}-33.5^{∘}and 67.5^{∘}-70.5^{∘}

J/AZh/77/96/  table1.dat  (35Kb)

Boyarchuk A.A., Antipova L.I., Boyarchuk M.E., Savanov I.S. Astron. Zh. 77, 96 (2000) Analysis of elemental abundances in the atmospheres of three Hyades red giants.

J/AZh/77/407/  table1.dat table3.dat table5.dat  (18Kb)

Gorshkov A.G., Konnikova V.K., Mingaliev M.G. Astron. Zh. 77, 407 (2000) Analysis of instantaneous spectra of a complete sample of flat-spectrum radio sources.

J/AZh/77/420/  table2.dat  (60Kb)

Rostopchina, A.N., Grinin, V.P., Shakhovskoy, D.N., The P.S., Minikulov N.Kh. Astron. Zh. 77, 420 (2000) Photopolarimetric activity of the Ae Herbig star SV Cep.

J/AZh/77/499/  table2.dat  (14Kb)

Malofeev V.M., Malov O.I., Shchegoleva N.V. Astron. Zh. 77, 499 (2000) Flux densities of 235 pulsars at 102.5 MHz

J/AZh/77/569/  table1.dat table2.dat  (25Kb)

Karachentseva V.E., Karachentsev I.D. Astron. Zh. 77, 569 (2000) Southern isolated triple systems of galaxies.

J/AZh/77/619/  table1.dat table2.dat table5.dat  (19Kb)

Knyazeva L.N., Kharitonov A.V. Astron. Zh. 77, 619 (2000) The normal energy distributions in stellar spectra: giants and supergiants.

J/AZh/78/60/  table1.dat  (22Kb)

Rostopchina A.N., Grinin V.P., Shakhovskoy D.N. Astron. Zh. 78, 60 (2001) Photometry and polarimetry of the classical Herbig Ae star VV Ser.

J/AZh/78/408/  table.dat  (43Kb)

Karitskaya E.A., Voloshina I.B., Goransky V.P., Grankin K.N., Dzhaniashvili I.B, Ezhkova O.V., Kochiashvili N.T., Kumsiashvili M.I., Kusakin A.V., Lyutyi V.M., Mel’Nikov S.Yu., Metlova N.V. Astron. Zh. 78, 408 (2001) Coordinated observations of Cyg X-1 (V1357) from 1994-1998 in the Commonwealth of Independent States.

J/AZh/78/675/  catalog.dat  (707Kb)

Bukhmastova Yu.L. Astron. Zh. 78, 675 (2001) Properties of quasar-galaxy associations and gravitational mesolensing by halo objects.

J/AZh/78/807/  table2.dat  (16Kb)

Antipova L.I., Boyarchuk A.A. Astron. Zh. 78, 807 (2001) The chemical composition of the red giant eta Ser.

J/AZh/78/922/  table3.dat table4.dat  (13Kb)

Trushkin S.A., Bursov N.N., Smirnova Yu.V. Astron. Zh. 78, 922 (2001) Six-day modulation of the quiescent radio emission of SS 433

J/AZh/78/963/  catalog.dat  (15Kb)

Chernin A.D., Kravtsova A.S., Zasov A.V., Arkhipova V.P. Astron. Zh. 78, 963 (2001) Galaxies with rows.

J/AZh/78/1104/  stars.dat table.dat  (16Kb)

Kovaleva D.A. Astron. Zh. 78, 1104 (2001) Age and metallicity estimates of intermediate mass stars in eclipsing binaries.

J/AZh/78/1135/  table1.dat table2.dat table3.dat table4.dat  (29Kb)

Tereshchenko V.M. Astron. Zh. 78, 1135 (2001) Intermediate-brightness spectrophotometric standards. Standards near +40^{∘}declination.

J/AZh/79/249/  table1.dat table2.dat  (33Kb)

Tereshchenko V.M. Astron. Zh. 79, 249 (2002) Intermediate-brightness equatorial spectrophotometric standards.

J/AZh/79/259/  objects.dat table1.dat table2.dat  (17Kb)

Kovaleva D.A. Astron. Zh. 79, 259 (2002) Studies of Intermediate-Mass Stellar Models using Eclipsing Binaries

J/AZh/79/309/  table1.dat table2.dat table3.dat  (20Kb)

Barsukova E.A., Borisov N.V., Goranskii V.P., Lyutyi V.M., Metlova N.V. Astron. Zh. 79, 309 (2002) A photometric and spectroscopic study of the CI Cam

J/AZh/79/328/  table1.dat table2.dat  (11Kb)

Val’tts I,V., Lyubchenko S.Yu. Astron. Zh. 79, 328 (2002) A study of class II methanol maser condensations in the star-forming regions W48

J/AZh/79/501/  table1.dat table2.dat  (14Kb)

Kuz’min A.D. Astron. Zh. 79, 501 (2002) A comparative analysis of the integrated radio luminosities of normal and millisecond pulsars.

J/AZh/79/510/  table1.dat  (11Kb)

Borkova T.V., Marsakov V.A. Astron. Zh. 79, 510 (2002) Subsystems of RR Lyrae variable Stars in our Galaxy.

J/AZh/79/526/  stars.dat table1.dat table2.dat  (13Kb)

Tarasova T.N.

Astron. Zh. 79, 526 (2002) Search for the magnetic fields in late-type giants.

J/AZh/79/542/  table1.dat table2.dat  (12Kb)

Bondar N.I. Astron. Zh. 79, 542 (2002) Variability of the photospheric radiation of active K-M dwarfs and their X-ray luminosities.

J/AZh/79/589/  table2.dat  (8Kb)

Verkhodanov O.V., Kopylov A.I., Pariiskii Yu.N., Soboleva N.S., Temirova A.V., Zhelenkova O.P. Astron. Zh. 79, 589 (2002) Color redshifts and the age of the stellar population of distant RC radio galaxies

J/AZh/79/610/  table2.dat table3.dat  (19Kb)

Alakoz A.V., Kalenskii S.V., Promislov V.G., Johansson L.E.B., Winnberg A. Astron. Zh. 79, 610 (2002) Parameters of warm molecular clouds from methyl acetylene observations

J/AZh/79/675/  table1.dat  (13Kb)

Babadzhanyants M.K., Belokon’ E.T. Astron. Zh. 79, 675 (2002) Optical monitoring of the BL Lac object ON 231 in 1971-1990.

J/AZh/79/738/  table1.dat table4.dat  (18Kb)

Tutukov A.V., Yungelson L.R. Astron. Zh. 79, 738 (2002) A model for the population of binary stars in the Galaxy.

J/AZh/79/841/  table1.dat  (8Kb)

Rachkovskaya T.M. Astron. Zh. 79, 841 (2002) Abundances of carbon, nitrogen, oxygen, and other elements in the atmosphere of the giant 20 Peg

J/AZh/79/867/  position.dat refs.dat table1.dat table2.dat table3.dat  (25Kb)

Holovatyy V.V., Melekh B.Ya. Astron. Zh. 79, 867 (2002) The chemical composition of HII regions in blue compact dwarf galaxies.

J/AZh/79/909/  table2.dat  (44Kb)

Boyarchuk A.A., Pakhomov Yu.V., Antipova L.I., Boyarchuk M.E. Astron. Zh. 79, 909 (2002) Analysis of chemical abundances in the atmospheres of moderate barium stars.

J/AZh/79/932/  table1.dat table2.dat  (15Kb)

Smirnov Yu.M. Astron. Zh. 79, 932 (2002) On the atomic constants of TaI.

J/AZh/79/963/  table1.dat  (9Kb)

Lipovka A.A., Lipovka N.M. Astron. Zh. 79, 963 (2002) The nature of active radio galaxies in the cluster A569

J/AZh/79/1087/  table1.dat table2.dat  (20Kb)

Sheffer E.K., Voloshina I.B., Goransky V.P., Lyuty V.M. Astron. Zh., 79, 1087 (2002) Optical features of the accretion disk in the HZ Her/Her X-1 system during the anomalous low state in 1999.

J/AZh/79/1118/  table1.dat  (8Kb)

Taranova O.G., Shenavrin V.I. Astron. Zh. 79, 1118 (2002) Infrared photometry of the unique object FG Sge in 1985-2001.

J/AZh/80/8/  catalog.dat  (26Kb)

Loktin A.V., Beshenov G.V. Astron. Zh. 80, 8 (2003) Proper motions of open star clusters and the rotation rate of the Galaxy.

J/AZh/80/266/  table.dat  (8Kb)

Smirnov Yu.M. Astron. Zh. 80, 266 (2003) On the atomic constants of ReI.

J/AZh/80/291/  catalog.dat  (48Kb)

Kharchenko N.V., Pakulyak L.K., Piskunov A.E. Astron. Zh. 80, 291 (2003) Subsystem of open clusters in the post-Hipparcos era: cluster structural parameters and proper motions.

J/AZh/80/458/  objects.dat table1.dat  (20Kb)

Mishenina T.V., Kovtyukh V.V., Korotin S.A., Soubiran C. Astron. Zh. 80, 458 (2003) Sodium abundances in in stellar atmospheres with differing metallicities.

J/AZh/80/535/  table.dat  (7Kb)

Raguzova N.V., Shugarov S.Yu., Ketsaris N.A. Astron. Zh. 80, 535 (2003) Observations and optical light-curve analysis for the precataclysmic variable BE Ursae Majoris.

J/AZh/80/704/  table.dat table3.dat  (60Kb)

Antipova L.I., Boyarchuk A.A., Pakhomov Yu.V., Panchuk V.E. Astron. Zh. 80, 704 (2003) Studies of classical barium stars.

J/AZh/80/978/  table1.dat table2.dat table3.dat  (23Kb)

Gorshkov A.G., Konnikova V.K., Mingaliev M.G. Astron. Zh. 80, 978 (2003) Spectra, optical identifications, and statistics of a complete sample of radio sources at declinations 10-12.5^{∘}.

J/AZh/81/33/  table.dat  (7Kb)

Smirnov Yu.M. Astron. Zh. 81, 33 (2004) Excitation of PtI transitions ending at the levels of the main term.

J/AZh/81/209/  stars.dat table1.dat  (12Kb)

Mashonkina L.I., Kamaeva L.A., Samotoev V.A., Sakhibullin N.A. Astron. Zh. 81, 209 (2004) Neutron capture elements in halo, thick disk and thin disk stars: neodymium.

J/AZh/81/333/  table1.dat table2.dat  (28Kb)

Klochkova V.G., Chentsov E.L. Astron. Zh. 81, 333 (2004) Kinematics of the envelope of the post-AGB star V510 Pup - nucleus of a future planetary nebula.

J/AZh/81/541/  table2.dat table3.dat table6.dat table7.dat table8.dat  (20Kb)

Galeev A.I., Bikmaev I.F., Musaev F.A., Galazutdinov G.A. Astron. Zh. 81, 541 (2004) Chemical Composition of 15 Photometric Analogues of the Sun.

J/AZh/81/658/  table2.dat  (65Kb)

Antipova L.I., Boyarchuk A.A., Pakhomov Yu.V., Panchuk V.E. Astron. Zh. 81, 658 (2004) Analysis of Atmospheric Abundances in Classical Barium Stars.

J/AZh/81/925/  table2.dat  (15Kb)

Bychkova V. S., Kardashev N.S., Vlasyuk V.V., Spiridonova O.I. Astron. Zh. 81, 925 (2004) The Rapid Variability of Blazars at Optical Wavelengths.

J/AZh/81/1032/  catalog.dat  (32Kb)

Titov O.A. Astron. Zh. 81, 1032 (2004) Construction of a Celestial Coordinate Reference Frame from VLBI Data.

J/AZh/81/1104/  table1.dat  (15Kb)

Klochkova V.G., Chentsov E.L. Astron. Zh. 81, 1104 (2004) The optical spectrum of the LBV candidate in the Cyg OB2 association.

J/AZh/82/231/  refs.dat stars.dat table1.dat  (15Kb)

Orlov V.V., Zhuchkov R.Ya. Astron. Zh. 82, 231 (2005) Analysis of the dynamic stability of selected multiple stars with weak hierarchy

J/AZh/82/420/  table1.dat table2.dat table3.dat  (15Kb)

Afanasiev V.L., Dodonov S.N., Moiseev A.V., Gorshkov A.G., Konnikova V.K., Mingaliev M.G.

Astron. Zh. 82, 420 (2005) Spectral studies with the Special Astrophysical Observatory 6m and RATAN-600 Telescopes.

J/AZh/82/437/  objects.dat refs.dat table1.dat table2.dat table3.dat table4.dat table5.dat  (39Kb)

Holovatyy V.V., Havrilova N.V. Astron. Zh. 82, 437 (2005) The chemical composition of galactic planetary nebulae with regard to inhomogeneity in the gas density in their envelopes.

J/AZh/82/453/  catalog.dat refs.dat table1.dat table3.dat  (378Kb)

Borkova T.V., Marsakov V.A. Astron. Zh. 82, 453 (2005) Compiled catalogue of spectroscopic abundances of different chemical element in stars with accurate parallaxes. I. Magnesium.

J/AZh/82/601/  table2.dat  (45Kb)

Antipova L.I., Boyarchuk A.A., Pakhomov Yu.V., Yushkin M.V. Astron. Zh. 82, 601 (2005) Analysis of the Na, Mg, Al, and Si Abundances in the Atmospheres of Red Giants of Different Spectral Subgroups.

J/AZh/83/241/  table1.dat table2.dat  (15Kb)

Gorshkov A.G., Konnikova V.K., Mingaliev M.G. Astron. Zh. 83, 241 (2006) Instantaneous spectra for a flux-density-complete sample of radio sources in the declination zone 74-75 deg.

J/AZh/83/265/  table3.dat  (22Kb)

Klochkova V.G., Panchuk V.E., Tavolganskaya N.S., Zhao G. Astron. Zh. 83, 265 (2006) Detection of spectral variability of the optical component of the IR-source IRAS 20508+2011.

J/AZh/83/542/  table2.dat table3.dat  (12Kb)

Malov I.F., Malov O.I. Astron. Zh. 83, 542 (2006) Integrated Radio Luminosities of Pulsars.

J/AZh/83/821/  osaca.dat  (8.7 Mb)

Bobylev V.V., Goncharov G.A., Bajkova A.T. Astron. Zh. 83, 821 (2006) The OSACA Database and a Kinematic Analysis of Stars in the Solar Neighborhood.

J/AZh/84/147/  v838mon.dat v4332sgr.dat  (52Kb)

Goranskii V.P., Barsukova E.A. Astron. Zh. 84, 147 (2007) Comparative spectral analysis of the peculiar red novae V838 Mon and V4332 Sgr in quiescence after their outbursts.

J/AZh/86/158/  objects.dat table1.dat  (23Kb)

Kovtyukh V.V., Mishenina T.V., Gorbaneva T.I., Bienayme O., Soubiran C., Kantcen L.E. Astron. Zh., 83, 158 (2006) Determination of high precision effective temperatures for giants based on spectroscopic criteria.

J/AcA/46/389/  table2.dat table3.dat table4.dat table5.dat table6.dat tables.tex  (200Kb)

Olech A. Acta Astron., 46, 389 (1996) The Miscellaneous Stars in the OGLE Catalog of Periodic Variable Stars in the Galactic Bulge

J/AcA/48/35/  table1.dat table2.dat  (18Kb)

Pojmanski G. Acta Astron., 48, 35 (1998) The All Sky Automated Survey. The catalogue of the periodic variables stars in the Selected Fields.

J/AcA/48/113/  table1.dat table2.dat table3.dat  (28Kb)

Udalski A. Acta Astron., 48, 113 (1998) The Optical Gravitational Lensing Experiment. The distance scale: Galactic Bulge - LMC - SMC.

J/AcA/48/147/ mr17079.fit mr17080.fit mr17081.fit mr17083.fit mr17084.fit mr17085.fit mr17086.fit mr17087.fit mr17088.fit mr17404.fit mr17417.fit smc sc1.dat smc sc2.dat smc sc3.dat smc sc4.dat smc sc5.dat smc sc6.dat smc sc7.dat smc sc8.dat smc sc9.dat smc sc10.dat smc sc11.dat (623.9 Mb)

Udalski A., Szymanski M., Kubiak M., Pietrzynski G., Wozniak P., Zebrun K. Acta Astron., 48, 147 (1998) The Optical Gravitational Lensing Experiment. BVI maps of dense stellar regions. I. The Small Magellanic Cloud.

J/AcA/48/175/  phot.dat table2.dat table3.dat  (2.3 Mb)

Pietrzynski G., Udalski A., Kubiak M., Szymanski M., Wozniak P., Zebrun K. Acta Astron., 48, 175 (1998) Optical Gravitational Lensing Experiment. The catalog of clusters in the Small Magellanic Cloud.

J/AcA/48/289/  imag.dat table2.dat vmag.dat  (735Kb)

Zebrun K. Acta Astron., 48, 289 (1998 Search for long-period and non-periodic variable stars in the OGLE-I databases

J/AcA/48/455/  m31gc.dat m33gc.dat  (19Kb)

Mochejska B.J., Kaluzny J., Krockenberger M., Sasselov D.D., Stanek K.Z. Acta Astron., 48, 455 (1998) Identification and photometry of globular clusters in M31 and M33 galaxies

J/AcA/49/59/  notes.dat table1.dat table2.dat table3.dat  (31Kb)

Weitenbeck A.J. Acta Astron., 49, 59 (1999) Stars with ISM polarization observed with HPOL

J/AcA/49/223/  lmc sc.dat table1.dat table4.dat  (251Kb)

Udalski A., Soszynski I., Szymanski M., Kubiak M., Pietrzynski G., Wozniak P., Zebrun K. Acta Astron., 49, 223 (1999) The Optical Lensing Experiment. Cepheids in the Magellanic Clouds. IV. Catalog of Cepheids from the Large Magellanic Cloud.

J/AcA/49/437/  smc sc.dat table1.dat table4.dat  (370Kb)

Udalski A., Soszynski I., Szymanski M., Kubiak M., Pietrzynski G., Wozniak P., Zebrun K. Acta Astron., 49, 437 (1999) The Optical Gravitational Lensing Experiment. Cepheids in the Magellanic Clouds. V. Catalog of Cepheids from the Small Magellanic Cloud.

J/AcA/49/521/  phot.dat table3.dat  (8.5 Mb)

Pietrzynski G., Udalski A., Kubiak M., Szymanski M., Wozniak P., Zebrun K. Acta Astron., 49, 521 (1999) The Optical Gravitational Lensing Experiment. Catalog of star clusters from the Large Magellanic Cloud.

J/AcA/49/543/  table1.dat table2.dat  (30Kb)

Pietrzynski G., Udalski A. Acta Astron., 49, 543 (1999) The Optical Gravitational Lensing Experiment. Cepheids in star clusters from the Magellanic Clouds.

J/AcA/50/1/  lc/* table1.dat table2.dat  (587Kb)

Udalski A., Zebrun K., Szymanski M., Kubiak M., Pietrzynski G., Soszynski I., Wozniak P. Acta Astron., 50, 1-65 (2000) The Optical Gravitational Lensing Experiment. Catalog of microlensing events in the Galactic Bulge.

J/AcA/50/177/  lcmisc.dat lcper.dat  (281Kb)

Pojmanski G. Acta Astron., 50, 177 (2000) The All Sky Automated Survey. Catalog of about 3800 variable stars.

J/AcA/50/221/  notes.dat table1.dat  (39Kb)

Bakos G.A., Benko J.M., Jurcsik J. Acta Astron., 50, 221 (2000) Identification and astrometry of variable stars in M 3.

J/AcA/50/337/  table1.dat  (34Kb)

Pietrzynski G., Udalski A. Acta Astron., 50, 337 (2000) The Optical Gravitational Lensing Experiment. Ages of about 600 star clusters from the LMC.

J/AcA/50/355/  table1.dat  (16Kb)

Pietrzynski G., Udalski A. Acta Astron., 50, 355 (2000) The Optical Gravitational Lensing Experiment. Multiple Cluster Candidates in the Large Magellanic Cloud.

J/AcA/50/421/  calib.dat catalog.dat  (744Kb)

Wozniak P.R. Acta Astron., 50, 421 (2000) Difference Image Analysis of the OGLE-II bulge data. I. The method.

J/AcA/51/175/  crossid.dat table1.dat tablesa.dat tablesb.dat  (99Kb)

Wozniak P.R., Udalski A., Szymanski M., Kubiak M., Pietrzynski G., Soszynski I., Zebrun K. Acta Astron., 51, 175 (2001) Difference Image Analysis of the OGLE-II bulge data. II. Microlensing events.

J/AcA/51/221/  iband/* ic1613.dat table1.dat vband/*  (1.9 Mb)

Udalski A., Wyrzykowski L., Pietrzynski G., Szewczyk O., Szymanski M., Kubiak M., Soszynski I., Zebrun K. Acta Astron., 51, 221 (2001) The optical gravitational lensing experiment. Cepheids in the galaxy IC 1613: no dependence of the periodluminosity relation on metallicity.

J/AcA/51/259/  ew.dat puls.dat  (230Kb)

Szymanski M., Kubiak M., Udalski A. Acta Astron., 51, 259 (2001) Contact binaries in OGLE-I database.

J/AcA/52/129/  catalog.dat  (31.4 Mb)

Wozniak P.R., Udalski A., Szymanski M., Kubiak M., Pietrzynski G., Soszinski I., Zebrun K. Acta Astron., 52, 129 (2002) Difference image analysis of the OGLE-II Bulge data. III. Catalog of 200000 candidate variable stars.

J/AcA/52/143/  lmc hpm.dat smc hpm.dat  (328Kb)

Soszynski I., Zebrun K., Udalski A., Wozniak P.R., Szymanski M., Kubiak M., Pietrzynski G., Szewczyk O., Wyrzykowski L. Acta Astron., 52, 143 (2002) The Optical Gravitational Lensing Experiment. Catalog of high proper motion stars towards the Magellanic Clouds

J/AcA/52/217/  fields.dat  (13Kb)

Udalski A., Szymanski M., Kubiak M., Pietrzynski G., Soszynski I., Wozniak P., Zebrun K., Szewczyk O., Wyrzykowski L. Acta Astron., 52, 217 (2002) The Optical Gravitational Lensing Experiment. BVI maps of dense stellar regions. III. The Galactic Bulge.

J/AcA/52/241/  table1.dat table2.dat  (16Kb)

Eyer L. Acta Astron., 52, 241 (2002) Search for QSO candidates in OGLE-II data.

J/AcA/53/1/  tablee.dat tableea.dat  (467Kb)

Wyrzykowski L., Udalski A., Kubiak M., Szymanski M., Zebrun K., Soszynski I., Wozniak P.R., Pietrzynski G., Szewczyk O. Acta Astron., 53, 1 (2003) The Optical Gravitational Lensing Experiment. Eclipsing binary stars in the Large Magellanic Cloud.

J/AcA/53/27/  idata/* tables.dat vdata/*  (9.5 Mb)

Pigulski A., Kolaczkowski Z., Kopacki G. Acta Astron., 53, 27 (2003) Reanalysis of the OGLE-I observations with the image substraction method. I. Galactic Bar fields MM1-A, MM1-B, MM7-A, and MM7-B.

J/AcA/53/93/  catalog.dat dmod.dat non rad.dat other.dat overlap.dat rrclust.dat  (1.3 Mb)

Soszynski I., Udalski A., Szymanski M., Kubiak M., Pietrzynski G., Wozniak P., Zebrun K., Szewczyk O., Wyrzykowski L. Acta Astron., 53, 93 (2003) The Optical Gravitanional Lensing Experiment. Catalog of RR Lyr stars in the Large Magellanic Cloud.

J/AcA/54/87/  table1.dat table2.dat table3.dat table4.dat table5.dat tables.ps tables.tex  (73Kb)

Weitenbeck A.J. Acta Astron., 54, 87-102 (2004) Stars with ISM polarization observed with HPOL. II.

J/AcA/54/207/  table1.dat  (244Kb)

Kreiner J.M. Acta Astron., 54, 207-210 (2004) Up-to-date linear elements of eclipsing binaries.

J/AcA/54/347/  ecl.dat ell.dat  (207Kb)

Soszinski I., Udalski A., Kubiak M., Szymanski M.K., Pietrzynski G., Zebrun K., Szewczyk O., Wyrzykowski L., Dziembowski W.A. Acta Astron., 54, 347-362 (2004) The optical gravitational lensing experiment. Ellipsoidal variability of red giants in the Large Magellanic Cloud.

J/AcA/56/253/ lc/* table1.dat table2.dat table3.dat (4.1 Mb) Kubiak M., Udalski A., Szymanski M.K. Acta Astron., 56, 253-265 (2006) Period changes of contact binaries in OGLE database.

J/ApJ/97/381 See III/13/ J/ApJ/104/234 See III/13/ J/ApJ/109/391 See II/221A/ J/ApJ/109/426 See II/221A/ J/ApJ/110/478 See II/221A/ J/ApJ/112/90 See II/221A/ J/ApJ/116/117 See III/13/ J/ApJ/118/77 See II/221A/ J/ApJ/130/178 See II/221A/ J/ApJ/143/306 See II/63A/ J/ApJ/157/327 See III/11/ J/ApJ/157/1279 See II/47/ J/ApJ/165/561 See II/45/ J/ApJ/170/L89 See II/56/ J/ApJ/252/102/ table1.dat table2.dat table3.dat table5.dat table6.dat table7.dat (31Kb)

Condon J.J., Condon M.A., Gisler G., Puschell J.J.

Astrophys. J. 252, 102 (1982)

Strong radio sources in bright spiral galaxies. II. Rapid star formation and galaxy-galaxy interactions. J/ApJ/255/382 See VII/85A/ J/ApJ/257/423 See VII/86A/ J/ApJ/261/1 See VII/48/ J/ApJ/261/412 See VII/171/ J/ApJ/263/777 See II/82/ J/ApJ/269/107 See VII/43/ J/ApJ/291/260 See V/65/ J/ApJ/297/582/ table2 table3 table4 (22Kb)

Da Costa G.S., Mould J.R., Crawford M.D.

Astrophys. J. 297, 582 (1985)

The age of the LMC globular cluster NGC 2213 J/ApJ/298/544/ table1 table2 (26Kb)

Stryker L.L., Da Costa G.S., Mould J.R.

Astrophys. J. 298, 544 (1985)

The main-sequence turnoff of the old SMC globular cluster NGC 121 J/ApJ/300/779/ notes.dat revised.dat table1.dat (38Kb)

Stephenson C.B.

Astrophys. J. 300, 779 (1986)

New H-alpha emission stars found above 10^{∘}galactic latitude. J/ApJ/314/493/ table1.dat table2.dat (25Kb)

Kirshner R.P., Oemler A.Jr, Schechter P.L., Shectman S.A.

Astrophys. J. 314, 493 (1987) A survey of the Bootes Void.

J/ApJ/315/687/  gstars 1.dat gstars 2.dat gstars 3.dat  (27Kb)

Maggio A., Sciortino S., Vaiana G.S., Majer P., Bookbinder J., Golub L., Harnden F.R.Jr, Rosner R. Astrophys. J. 315, 687 (1987) Einstein Observatory survey of X-ray emission from solar-type stars: the late F and G dwarf stars.

J/ApJ/317/102/  table2.dat table3.dat  (19Kb)

Salpeter E.E., Dickey J.M. Astrophys. J. 317, 102 (1987) 1.4 GHz continuum sources in the Cancer cluster.

J/ApJ/322/706 See VIII/39/

J/ApJ/324/172/  table1  (57Kb)

Hodge P., Lee M.G., Mateo M. Astrophys. J. 324, 172 (1988) UBVR CCD photometry of stars in M 31. I. Baade’s field IV

J/ApJ/325/798/  hyades 1.dat hyades 2.dat  (51Kb)

Micela G., Sciortino S., Vaiana G.S., Schmitt J.H.M.M., Stern R.A., Harnden F.R.Jr, Rosner R. Astrophys. J. 325, 798 (1988) The Einstein Observatory survey of stars in the Hyades cluster region.

J/ApJ/328/315 See III/116/

J/ApJ/336/780/  table3  (28Kb)

Nemec J.M., Cohen J.G. Astrophys. J. 336, 780 (1989) Blue straggler stars in the globular cluster NGC 5053

J/ApJ/339/53/  table1.dat table2.dat table6.dat  (30Kb)

Ciardullo R., Jacoby G.H., Ford H.C., Neill J.D. Astrophys. J. 339, 53 (1989) Planetary nebulae as standard candles. II. The calibration in M 31 and its companions.

J/ApJ/339/126 See V/65/

J/ApJ/344/704/  table1.dat table2.dat  (13Kb)

Jacoby G.H., Ciardullo R., Ford H.C., Booth J. Astrophys. J. 344, 704 (1989) Planetary nebulae as standard candles. III. The distance to M 81.

J/ApJ/344/715/  tab2-4.dat table5.dat  (21Kb)

Ciardullo R., Jacoby G.H., Ford H.C. Astrophys. J. 344, 715 (1989) Planetary nebulae as standard candles. IV. A test in the Leo I group.

J/ApJ/347/251/  table1.dat table2.dat table3.dat  (21Kb)

Peterson Ruth C., Seitzer Patrick, Cudworth Kyle M. Astrophys. J. 347, 251 (1989) (From AAS CD-ROM series, Volume 4, 1995 Lee Brotzman [ADS] 06-May-1995) The Nonthermal Stellar Dynamics of the Globular Cluster M15

J/ApJ/348/98/  table3  (18Kb)

Reid N., Tinney C., Mould J. Astrophys. J. 348, 98 (1990) Luminous asymptotic giant branch stars in the Large Magellanic Cloud

J/ApJ/348/253/  giants.dat  (69Kb)

Maggio A., Vaiana G.S., Haisch B.M., Stern R.A., Bookbinder J., Harnden F.R.Jr, Rosner R. Astrophys. J. 348, 253 (1990) EINSTEIN Observatory magnitude-limited X-ray survey of late-type giant and supergiant stars.

J/ApJ/348/557/  pleiad 1.dat pleiad 2.dat refs.dat  (47Kb)

Micela G., Sciortino S., Vaiana G.S., Harnden F.R., Rosner R., Schmitt J.H.M.M. Astrophys. J. 348, 557 (1990) X-ray studies of coeval star samples. II. The Pleiades cluster as observed with the Einstein Observatory.

J/ApJ/351/492/  st uma 1.dat st uma 3.dat  (17Kb)

Schmitt J.H.M.M., Micela G., Sciortino S., Vaiana G.S., Harnden F.R., Rosner R. Astrophys. J. 351, 492 (1990) X-ray studies of coeval star samples. III. X-ray emission in the Ursa Major stream.

J/ApJ/353/494/  table2.dat table3.dat table4.dat table5.dat table6.dat  (38Kb)

Frogel J.A., Terndrup D.M., Blanco V.M., Whitford A.E. Astrophys. J. 353, 494 (1990) Galactic bulge M Giants. II. Content and structure of the bulge between b = -3 and -12

J/ApJ/356/332/  tab4-9.dat table10.dat  (29Kb)

Jacoby G.H., Ciardullo R., Ford H.C. Astrophys. J. 356, 332 (1990) Planetary nebulae as standard candles. V. The distance to the Virgo cluster.

J/ApJ/361/49 See VII/185/

J/ApJ/365/66/  table1 table2  (41Kb)

HUCHRA J.P., HENRY J.P., POSTMAN M., GELLER M.J. Astrophys. J. 365, 65 (1990) A deep Abell cluster redshift survey

J/ApJ/365/471/  table2.dat table3.dat  (10Kb)

Jacoby G.H., Walker A.R., Ciardullo R. Astrophys. J. 365, 471 (1990) Planetary nebulae as standard candles. VI. A test in the Magellanic Clouds.

J/ApJ/368/54/  table1  (15Kb)

DRESSLER A., FABER S.M., BURSTEIN D. Astrophys. J. 368, 54 (1991) New velocity dispersions and photometry for E and S0 galaxies in the great attractor

J/ApJ/374/440 See IX/22/

J/ApJ/378/65/  table1.dat table2.dat  (16Kb)

Condon J.J., Huang Z.-P., Yin Q.F., Thuan T.X. Astrophys. J. 378, 65 (1991) Compact starbursts in ultraluminous infrared galaxies.

J/ApJ/380/30 See VII/213/

J/ApJ/382/636/  table1.dat  (11Kb)

Itoh N., Kuwashima F., Ichihashi K., Mutoh H. Astrophys. J. 382, 636 (1991) (From AAS CD-ROM series, Volume 9, 1997 Lee E. Brotzman [ADS] 27-Aug-97) The Rosseland mean free-free Gaunt factor of the dense high-Temperature stellar plasma

J/ApJ/383/487/  table2.dat table3.dat table5.dat  (16Kb)

Ciardullo R., Jacoby G.H., Harris W.E. Astrophys. J. 383, 487 (1991) Planetary nebulae as standard candles. VII. A test versus Hubble type in the NGC 1023 group.

J/ApJ/384/404/  table1  (27Kb)

POSTMAN M., HUCHRA J. P., GELLER M. J. Astrophys. J. 384, 404 (1992) The distribution of nearby rich clusters of galaxies

J/ApJ/391/685 See VI/55/

J/ApJ/393/149/  table4.dat  (14Kb)

Gwinn C.R., Moran J.M., Reid M.J. Astrophys. J. 393, 149 (1992) Distance and kinematics of the W49N H2O maser outflow

J/ApJ/395/L1 See VI/104/

J/ApJ/397/520/  table1 table2 table3  (29Kb)

WILKING B.A., THOMAS P.G., LADA C.J., MEYER M.R., YOUNG E.T. Astrophys. J. 397, 520 (1992) IRAS observations of young stellar objects in the Corona Australis dark cloud

J/ApJ/399/114 See VI/56/

J/ApJ/399/353 See VII/213/

J/ApJ/400/699 See VI/62/

J/ApJ/405/94/  table3.dat  (21Kb)

Wise M.W., O’Connell R.W., Bregman J.N., Roberts M.S. Astrophys. J. 405, 94 (1993) Far-infrared emission from the intracluster medium.

J/ApJ/405/498/  table1a table1b  (16Kb)

WINDHORST R.A., FOMALONT E.B., PARTRIDGE R.B., LOWENTHAL J.D. Astrophys. J. 405, 498 (1993) Microjansky source counts and spectral indices at 8.44 GHz

J/ApJ/408/305/  table1  (27Kb)

LANDSMAN W.B. , SIMON T. Astrophys. J. 408, 305 (1993) A catalog of stellar Lyman-alpha fluxes

J/ApJ/408/484/  table1  (38Kb)

SCHMIDT E.G., CARRUTHERS G.R. Astrophys. J. 408, 484 (1993) Far-ultraviolet stellar photometry: a field in Monoceros

J/ApJ/409/28/  table1a.dat table1b.dat  (52Kb)

Maoz D., Bahcall J.N., Schneider D.P., Bahcall N.A., Djorgovski S., Gould A., Kirhakos S., Doxsey R., Meylan G., Yanny B. Astrophys. J. 409, 28 (1993) The Hubble Space Telescope Snapshot Survey. IV A Summary of the Search for Gravitationally Lensed Quasars

J/ApJ/411/188/  tables  (40Kb)

ALLEN L.E., KLEINMANN S.G., WEINBERG M.D. Astrophys. J. 411, 188 (1993) IRAS variables as galactic structure tracers: classification of the bright variables.

J/ApJ/412/183/  table3 table4  (37Kb)

CLEMENT C.M., FERANCE S., SIMON N.R. Astrophys. J. 412, 183 (1993) (From AAS CD-ROM series, Volume 1, 1993) The RR Lyrae variables in the globular cluster M68

J/ApJ/413/453/  refs reftab1 table1  (70Kb)

NILSSON K., VALTONEN M.J., KOTILAINEN J., JAAKKOLA T. Astrophys. J. 413, 453 (1993) On the redshift-apparent size diagram of double radio sources

J/ApJ/413/604/  aastab.sty environ.tex table3 table3.tex table5 table5.tex  (80Kb)

HILL J.K., BOHLIN R.C., CHENG K.-P., FANELLI M.N., HINTZEN P., O’CONNELL R.W., ROBERTS M.S., SMITH A.M., SMITH E.P., STECHER T.P. Astrophys. J. 413, 604 (1993) (From AAS CD-ROM series, Volume 1, 1993) 30 Doradus: ultraviolet and optical stellar photometry

J/ApJ/414/279/  table1 table2 table3  (53Kb)

TINNEY C.G. Astrophys. J. 414, 279 (1993) The faintest stars: the luminosity and mass functions at the bottom of the main sequence.

J/ApJ/414/454/  table1.dat  (7Kb)

Ciardullo R., Jacoby G.H., Dejonghe H.B. Astrophys. J. 414, 454 (1993) The radial velocities of planetary nebulae in NGC 3379.

J/ApJ/414/463/  table2.dat table6.dat  (11Kb)

Hui X.-H., Ford H.C., Ciardullo R., Jacoby G.H. Astrophys. J. 414, 463 (1993) The planetary nebula system and dynamics of NGC 5128. I. Planetary nebulae as standard candles.

J/ApJ/414/846/  table1.dat  (36Kb)

Barbera M., Micela G., Sciortino S., Harden F.R. Jr, Rosner R. Astrophys. J. 414, 846 (1993) X-ray emission at the low-mass end: results from an extensive ”Einstein Observatory” survey

J/ApJ/416/36/  table1.dat table1a.dat  (27Kb)

Pinkney J., Rhee G., Burns J.O., Hill J.M., Oegerle W., Batuski D., Hintzen P. Astrophys. J. 416, 36 (1993) The Dynamics of the Galaxy Cluster Abell 2634

J/ApJ/416/62/  tab3-5.dat table7.dat  (19Kb)

McMillan R., Ciardullo R., Jacoby G.H. Astrophys. J. 416, 62 (1993)

Planetary nebulae as standard candles. IX. The distance to the Fornax cluster.

J/ApJ/416/276/  b10 0.dat b10 5.dat b11 0.dat b11 5.dat b12 0.dat b12 5.dat b13 0.dat  (26Kb)

Rognvaldsson O.E., Fushiki I., Gudmundsson E.H., Pethick C.J., Yngvason J. Astrophys. J., 416, 276 (1993) Thomas-Fermi calculations of atoms and matter in magnetic neutron stars: effects of higher Landau bands

J/ApJ/416/787/  table1 table2  (36Kb)

SOON W.H., ZHANG Q., BALIUNAS S.L., KURUCZ R.L. Astrophys. J. 416, 787 (1993) The Mount Wilson Observatory Metallicity index, Crv: comparison with other photometric systems

J/ApJ/417/209/  table2.dat table2a.dat table4.dat table5.dat  (18Kb)

Jacoby G.H., Kaler J.B. Astrophys. J. 417, 209 (1993) Improved observations of faint planetary nebulae in the Magellanic Clouds.

J/ApJ/417/502 See VI/76/

J/ApJ/419/596 See VI/65/

J/ApJ/420/612/  table2.dat table3.dat table4.dat  (22Kb)

Peterson R.C., Cudworth K.M. Astrophys. J. 420, 612 (1994) (From AAS CD-ROM series, Volume 4, 1995 Lee Brotzman [ADS] 06-May-1995) Proper motions and radial velocities in the globular cluster M22 and the cluster distance

J/ApJ/424/599/  table5 table6 table8  (19Kb)

Martin P., Roy J-R. Astrophys. J. 424, 599 (1994) The influence of bars on the chemical composition of spiral galaxies

J/ApJ/425/122/  aastab.sty environ.tex table2.dat table2.tex  (230Kb)

HILL J.K., ISENSEE J.E., CORNETT R.H., BOHLIN R.C., O’CONNELL R.W., ROBERTS M.S., SMITH A.M., STECHER T.P. Astrophys. J. 425, 122 (1994) (From AAS CD-ROM series, Volume 2, 1994) Initial mass functions from ultraviolet stellar photometry: a comparison of Lucke and Hodge OB Associations near 30 Doradus with the nearby field

J/ApJ/427/125/  table1 table2 table3  (14Kb)

Lister M.L., Hutchings J.B., Gower A.C. Astrophys. J. 427, 125 (1994) Orientation modeling of radio galaxy and quasar properties: evidence for a unified model.

J/ApJ/427/684 See VII/213/

J/ApJ/428/693/  table2.dat table3a.dat table3b.dat  (17Kb)

Williams J.P., De Geus E.J., Blitz L. Astrophys. J. 428, 693 (1994) Determining structure in molecular clouds.

J/ApJ/429/253/  table1  (15Kb)

Gwinn C.R. Astrophys. J. 429, 253 (1994) Physical structure of H2O masers in W49N

J/ApJ/430/467/  notes.dat refs.dat table1.dat  (28Kb)

Vermeulen R.C., Cohen M.H. Astrophys. J. 430, 467 (1994) Superluminal motion statistics and cosmology.

J/ApJ/434/54/  table1a table1b table1c  (26Kb)

Thompson R.J., Cordova F.A. Astrophys. J. 434, 54 (1994) A statistical analysis of the broadband 0.1 to 3.5keV spectral properties of X-ray-selected active galactic nuclei.

J/ApJ/437/361/  table2.dat table3.dat table6.dat table7.dat table8.dat tables.tex  (114Kb)

Gagne M., Caillault J.-P. Astrophys. J. 437, 361 (1994) The complete Einstein Observatory X-ray survey of the Orion nebula region.

J/ApJ/438/181/  aastab.sty environ.tex table2.dat table2.tex  (22Kb)

Hill J.K., Cheng K.-P., Bohlin R.C., Cornett R.H., Hintzen P.M.N., Robert M.S., O’Connell R.W., Roberts M.S.,Smith A.M., Smith E.P., Stecher T.P. Astrophys. J. 438, 181 (1995) (From AAS CD-ROM series, Volume 4, 1995 Lee Brotzman [ADS] 06-May-1995) Ultraviolet Imaging Telescope and Optical Emission-Line Observations of H II Regions in M81

J/ApJ/438/813/  table1 table2 table3 table4 table6  (41Kb)

STROM K.M., KEPNER J., STROM S.E. Astrophys. J. 438, 813 (1995) The evolutionary status of the stellar population in the ρOphiuchi cloud core

J/ApJ/439/983/  optical radio  (22Kb)

Elias N.M. II, Quirrenbach A., Witzel A., Naundorf C.E., Wegner R., Guinan E.F., McCook G.P. Astrophys. J. 439, 983 (1995) (From AAS CD-ROM series, Volume 4, 1995 Lee Brotzman [ADS] 06-May-1995) Correlations between the flaring radio emission and starspot distribution of UX Arietis

J/ApJ/440/28/  table1.dat table3.dat table4.dat table5.dat table6.dat table7.dat  (50Kb)

Postman M., Lauer T.R Astrophys. J. 440, 28 (1995) Brightest cluster galaxies as standard candles.

J/ApJ/441/307/  table3  (39Kb)

Anderson M.C., Rudnick L. Astrophys. J. 441, 307 (1995) The deceleration powering of synchrotron emission from ejecta components in supernova remnant Cassiopeia A

J/ApJ/442/523/  tablea1a.dat tablea1b.dat tablea2.dat tablea3.dat tablea4.dat  (55Kb)

Eskridge P.B., Fabbiano G., Kim D.-W. Astrophys. J. 442, 523 (1995) (From AAS CD-ROM series, Volume 4, 1995 Lee Brotzman [ADS] 06-May-1995) A Multiparametric Analysis of the Einstein Sample of Early-Type Galaxies. II. Galaxy Formation History and Properties of the Interstellar Medium

J/ApJ/443/124/  table1.dat table2.dat table3.dat table5.dat  (66Kb)

Peterson R.C., Rees R.F., Cudworth K.M. Astrophys. J. 443, 124 (1995) (From AAS CD-ROM series, Volume 4, 1995 Lee Brotzman [ADS] 06-May-1995) Radial Velocities of Stars in the Globular Cluster M4 and the Cluster Distance

J/ApJ/445/62/  table2.dat  (24Kb)

Hooper E.J., Impey C.D., Foltz C.B., Hewett P.C. Astrophys. J. 445, 62 (1995) Radio properties of optically selected quasars.

J/ApJ/445/280/  table2 table2.tex table3 table3.tex table6 table6.tex  (159Kb)

Gagne M., Caillault J.-P., Stauffer J.R. Astrophys. J. 445, 280 (1995) Deep ROSAT HRI observations of the Orion nebula region.

J/ApJ/446/12 See VII/198/

J/ApJ/446/115/  table4.dat table6.dat  (233Kb)

Vrba F.J., Hartmann D.H., Jennings M. C. Astrophys. J. 446, 115 (1995) Deep optical counterpart searches of gamma-ray burst localizations

J/ApJ/446/300/  fig1.dat fig2.dat fig3.dat fig4.dat fig5.dat  (267Kb)

Weaver W.B., Torres-Dodgen A.V. Astrophys. J. 446, 300 (1995) (From AAS CD-ROM series, Volume 5, 1995 Lee Brotzman [ADS] 07-Oct-95) Neural network classification of the near-infrared spectra of A-type stars

J/ApJ/446/622/  lh52 bv.dat lh53 bv.dat table2.dat table3.dat  (230Kb)

Hill R.S., Cheng K.-P., Bohlin R.C., OConnell R.W., Roberts M.S., Smith A.M., Stecher T.P. Astrophys. J. 446, 622 (1995) (From AAS CD-ROM series, Volume 5, 1995 Lee Brotzman [ADS] 10-Oct-95) UIT and optical imagery of Large Magellanic cloud associations LH 52 and LH 53: ages and initial mass function slopes

J/ApJ/448/179/  table2.dat  (357Kb)

Hunter D.A., Shaya E.J., Holtzman J.A., Light R.M., Earl J., Oneil Jr. Astrophys. J. 448, 179 (1995) (From AAS CD-ROM series, Volume 5, 1995 Lee Brotzman [ADS] 10-Oct-95)

The intermediate stellar mass population in R136 determined from Hubble Space Telescope planetary camera 2 images

J/ApJ/448/683/  table1 table2 table3  (60Kb)

Stern R.A., Schmitt J.H.M.M., Kahabka P.T. Astrophys. J. 448, 683 (1995) ROSAT all-sky survey observations of the Hyades cluster

J/ApJ/449/164/  idata vdata  (102Kb)

Sebo P.R., Wood K.M. Astrophys. J. 449, 164 (1995) (From AAS CD-ROM series, Volume 5, 1995 Lee Brotzman [ADS] 10-Oct-95) Variable Stars in Magellanic Cloud Clusters. II. NGC 1850

J/ApJ/449/527/  table2 table3  (41Kb)

Lu N.Y., Freudling W. Astrophys. J. 449, 527 (1995) Large-scale structures in the zone of avoidance: the galactic anticenter region

J/ApJ/449/592/  table5  (29Kb)

Hui X., Ford H.C., Freeman K.C., Dopita M.A. Astrophys. J. 449, 592 (1995) The planetary nebula system and dynamics of NGC 5128. III. Kinematics and halo mass distributions

J/ApJ/450/51/  table1 tables.tex  (216Kb)

Green P.J., Schartel N., Anderson S.F., Hewett P.C., Foltz C.B., Brinkmann W., Fink H., Trumper J., Margon B. Astrophys. J. 450, 51 (1995) (From AAS CD-ROM series, Volume 4, 1995 Lee Brotzman [ADS] 10-Oct-95) The soft X-ray properties of a large optical QSO sample: ROSAT observations of the Large Bright Quasar Survey

J/ApJ/452/210/  table1a.dat table1b.dat table2.dat  (36Kb)

Oey M.S., Massey P. Astrophys. J. 452, 210 (1995) Triggered star formation and the dynamics of a superbubble in the LMC: the OB association LH 47/48 in DEM 152

J/ApJ/452/238/  table2.dat  (45Kb)

Hunter H.A., Thronson H.A. Astrophys. J. 452, 238 (1995) (From AAS CD-ROM series, Volume 5, 1995 Lee Brotzman [ADS] 11-Oct-95) The Massive Stars of I Zw 18 as Seen in Hubble Space Telescope Images

J/ApJ/453/48/  table2  (18Kb)

Driver S.P., Windhorst R.A., Griffiths R.E. Astrophys. J. 453, 48 (1995) The contribution of late-type/irregulars to the faint galaxy counts from Hubble Space Telescope Medium Deep Survey images

J/ApJ/453/616/  table1 table2 table3.dat  (81Kb)

Spinoglio L., Malkan M.A., Rush B., Carrasco L., Recillas-Cruz E. Astrophys. J. 453, 616 (1995) (From AAS CD-ROM series, Volume 5, 1995 Lee Brotzman [ADS] 04-Nov-95) Multiwavelength energy distributions and bolometric luminosities of the 12 micron galaxy sample

J/ApJ/454/151/ berk 86.dat cep ob5.dat ic1805.dat ngc1893.dat ngc2244.dat ngc6823.dat ngc6871.dat ngc6913.dat ngc7235.dat ngc7380.dat (452Kb) Massey P., Johnson K.E., De Gioia-Eastwood K. Astrophys. J. 454, 151 (1995) (From AAS CD-ROM series, Volume 5, 1995 Lee Brotzman [ADS] 05-Nov-95) The Initial Mass Function and Massive Star Evolution in the OB Associations of the Northern Milky Way

J/ApJ/454/788/  table3  (117Kb)

Cote P., Welch D.L., Fischer P., Gebhardt K. Astrophys. J. 454, 788 (1995) (From AAS CD-ROM series, Volume 5, 1995 Lee Brotzman [ADS] 05-Nov-95) Dynamics of the Galactic globular cluster NGC 3201

J/ApJ/455/50 See VII/225A/

J/ApJ/455/60 See VII/225A/

J/ApJ/455/75 See VII/225A/

J/ApJ/455/88 See VII/225A/

J/ApJ/455/96 See VII/225A/

J/ApJ/456/174/  table2.dat  (3.1 Mb)

Hunter D.A., Baum W.A., O’Neil E.J.JR., Lynds R. Astrophys. J. 456, 174 (1996) (From AAS CD-ROM series, Volume 6, 1996 Lee Brotzman [ADS] 22-Apr-96) The intermediate stellar mass population in NGC 604 determined from Hubble Space Telescope Images

J/ApJ/456/234/  table1  (29Kb)

Anderson M.C., Rudnick L. Astrophys. J. 456, 234 (1996) Sites of relativistic particle acceleration in supernova remnant Cassiopeia A

J/ApJ/457/102/ q1206119.dat q1225017.dat q1315472.dat q1607183.dat q1623268.dat q1700643.dat q1946770.dat table2.tex (118Kb) Dobrzycki A., Bechtold J. Astrophys. J. 457, 102 (1996) (From AAS CD-ROM series, Volume 6, 1996 Lee Brotzman [ADS] 22-Apr-96) Simulation analysis of Ly-Alpha forest spectra. I. Empirical description at z 3

J/ApJ/457/460 See VII/198/

J/ApJ/458/455/  table3 table6 table7  (21Kb)

Ford H.C., Hui X., Ciardullo R., Jacoby G.H., Freeman K.C. Astrophys. J. 458, 455 (1996) The stellar halo of M104. I. A survey for planetary nebulae and the planetary nebula luminosity function distance

J/ApJ/459/110/  table9 table10  (81Kb)

Fisher D., Franx M., Illingworth G. Astrophys. J. 459, 110 (1996) (From AAS CD-ROM series, Volume 6, 1996 Lee Brotzman [ADS] 22-Apr-96) Line strengths and line strength gradients in S0 galaxies

J/ApJ/459/278/ hr3831b.dat hr3831h.dat hr3831i.dat hr3831j.dat hr3831k.dat hr3831r.dat hr3831v.dat hr3831y.dat (29Kb) Matthews J.M., Wehlau W.H., Rice J., Walker G.A.H. Astrophys. J. 459, 278 (1996) (From AAS CD-ROM series, Volume 6, 1996 Lee Brotzman [ADS] 22-Apr-96) Opening a new window on Ap star atmospheres: a T-Tauri relation for HR 3831 from its limb-darkened pulsation amplitudes

J/ApJ/459/606/  table1  (11Kb)

Schneider S.E., Buckley D. Astrophys. J. 459, 606 (1996) The distances of planetary nebulae and the galactic bulge

J/ApJ/459/686/  graphite.dat silicate.dat  (29Kb)

Dwek E., Smith R.K. Astrophys. J. 459, 686 (1996) Energy Deposition and Photoelectric Emission from the Interaction of 10 eV to 1 MeV Photons with Interstellar Dust Particles

J/ApJ/461/127/  twosigma.dat  (3.6 Mb)

Moran E.C., Helfand D.J., Becker R.H., White R.L. Astrophys. J. 461, 127 (1996) The Einstein Two-Sigma Catalog

J/ApJ/462/1/  table2 table3 table4 table6  (20Kb)

Jacoby G.H., Ciardullo R., Harris W.E. Astrophys. J. 462, 1 (1996) Planetary nebulae as standard candles. X. Tests in the Coma I region.

J/ApJ/462/241/  table1.dat table2.dat table3.dat  (40Kb)

Norris J.E., Freeman K.C., Mighell K.J. Astrophys. J. 462, 241N (1996) The giant branch of ωCentauri. V. The calcium abundance distribution.

J/ApJ/462/614/  table1 table1.tex table2 table2.tex  (20Kb)

Hall P.B., Osmer P.S., Green R.F., Porter A.C., Warren S.J. Astrophys. J. 462, 614 (1996) A deep multicolor survey. II. Initial spectroscopy and comparison with expected quasar number counts.

J/ApJ/463/26/  tableb1 tableb2  (45Kb)

Kelson D.D., Illingworth G.D., Freedman W.F., Graham J.A., Hill R., Madore B.F., Saha A., Stetson P.B., Kennicutt Jr.R.C., Mould J.R., Hughes S.M., Ferrarese L., Phelps R., Turner A., Cook K.H., Ford H., Hoessel J.G., Huchra J. Astrophys. J. 463, 26 (1996) (From AAS CD-ROM series, Volume 7, 1996 Lee Brotzman [ADS] 10-Oct-96)

The Extragalactic Distance Scale Key Project. III. The discovery of Cepheids and a new distance to M101 using the Hubble Space Telescope.

J/ApJ/463/205/  aaspp4.sty csh2o.tex csh2o.txt refs.dat table1.dat table2.dat table5.dat  (181Kb)

Anglada G., Estalella R., Pastor J., Rodriguez L.F., Haschick A.D. Astrophys. J. 463, 205 (1996) A CS and NH3 Survey of H2O Maser Emission

J/ApJ/464/79/  table1.dat  (13Kb)

Schade D., Lilly S.J., Le Fevre O., Hammer H., Crampton D. Astrophys. J. 464, 79 (1996) Canada-France Redshift Survey. XI. Morphology of high-redshift field galaxies from high-resolution groundbased imaging

J/ApJ/466/254/  table3.dat  (55Kb)

Brandl B., Sams B.J., Bertoldi F., Eckart A., Genzel R., Drapatz S., Hofmann R., Loewe M., Quirrenbach A. Astrophys. J. 466, 254 (1996) (From AAS CD-ROM series, Volume 7, 1996 Lee Brotzman [ADS] 22-Oct-96) Adaptive optics near-infrared imaging of R136 in 30 Doradus: the stellar population of a nearby starburst.

J/ApJ/466/415/  table.dat  (220Kb)

Mazeh T., Latham D.W., Stefnik R.P. Astrophys. J. 466, 415 (1996) (From AAS CD-ROM series, Volume 7, 1996 Lee Brotzman [ADS] 22-Oct-96) Spectroscopic orbits for three binaries with low-mass companions and the distribution of secondary masses near the substellar limit

J/ApJ/468/633/  table2.dat table3.dat  (2.9 Mb)

Hunter D.A., Baum W.A., O’Neil E.J.Jr, Lynds R. Astrophys. J. 468, 633 (1996) The intermediate stellar mass population in the M31 OB association NGC 206.

J/ApJ/469/629/  table1.dat  (38Kb)

Massey P., Bianchi L., Hutchings J.B., Stecher T.P. Astrophys. J. 469, 629 (1996) (From AAS CD-ROM series, Volume 7, 1996 Lee Brotzman [ADS] 22-Oct-96) The UV-brightest stars of M33 and its nucleus: discovery, photometry, and optical spectroscopy

J/ApJ/470/172 See VII/203/

J/ApJ/470/724/  table1.dat  (20Kb)

Mohr J.J., Geller M.J., Fabricant D.G., Wegner G., Thorstensen J., Richstone D.O. Astrophys. J. 470, 724 (1996) An optical and X-ray study of Abell 576, a galaxy cluster with a cold core.

J/ApJ/470/864/  notes.dat table1.dat table2.dat table4.dat  (44Kb)

Blum R.D., Sellgren K., Depoy D.L. Astrophys. J. 470, 864 (1996) JHKL photometry and the K-band luminosity function at the Galactic Center.

J/ApJ/471/L5/  table1  (12Kb)

Cohen J.G., Cowie L.L., Hogg D.W., Songaila A., Blandford R., Hu E.M., Shopbell P. Astrophys. J. 471, L5 (1996) Redshift clustering in the Hubble Deep Field.

J/ApJ/471/694/  table1.dat  (40Kb)

Abraham R.G., Smecker-Hane T.A., Hutchings J.B., Carlberg R.G., Yee H.K.C., Ellingson E., Morris S., Oke J.B., Rigler M. Astrophys. J. 471, 694 (1996) Galaxy evolution in Abell 2390.

J/ApJ/473/822/  table3  (164Kb)

Hoffman G.L., Dickey J.M., Lu N.Y., Fromhold-Treu R. Astrophys. J. 473, 822 (1996) A Minnesota automated plate scanner catalog of galaxies behind the Virgo cluster and toward its antipode

J/ApJ/475/445/  table4 table5  (19Kb)

Moustakas L.A., Davis M., Graham J.R., Silk J., Peterson B.A., Yoshii Y. Astrophys. J. 475, 445 (1997) Colors and K-band counts of extremely faint field galaxies.

J/ApJ/475/469/  fig5 fig8a fig8b  (988Kb)

Zheng W., Kriss G.A., Telfer R.C., Grimes J.P., Davidsen A.F. Astrophys. J. 475, 469 (1997) (From AAS CD-ROM series, Volume 8, 1997 Lee E. Brotzman [ADS] 06-May-

A Composite HST Spectrum of Quasars

J/ApJ/475/479 See VIII/48/

J/ApJ/476/311/  table2.dat table4.dat  (9Kb)

Kirkpatrick J.D., Beichman C.A., Skrutskie M.F. Astrophys. J. 476, 311K (1997) The coolest isolated M dwarf and other 2MASS discoveries.

J/ApJ/478/603/  table1 table2  (9Kb)

Bally J., Devine D., Alten V., Sutherland R.S. Astrophys. J. 478, 603 (1997) New Herbig-Haro flows in L1448 and L1455

J/ApJ/479/231/  tab2-4.dat  (14Kb)

Feldmeier J.J., Ciardullo R., Jacoby G.H. Astrophys. J. 479, 231 (1997) Planetary nebulae as standard candles. XI. Application to spiral galaxies

J/ApJ/479/427/  table3 table3.tex table6 table6.tex  (51Kb)

Gonzalez G., Lambert D.L., Giridhar S. Astrophys. J. 479, 427 (1997) (From AAS CD-ROM series, Volume 8, 1997 Lee E. Brotzman [ADS] 06-May-97) Abundance Analyses of the Field RV Tauri Variables: EP Lyrae, DY Orionis, AR Puppis, and R Sagittae

J/ApJ/480/568/  table4  (15Kb)

Steidel C.C., Dickinson M., Meyer D.M., Adelberger K.L., Sembach K.R. Astrophys. J. 480, 568 (1997) Quasar absorbing galaxies at z =1. I. Deep imaging and spectroscopy in the field of 3C 336.

J/ApJ/481/49/  table1.dat  (78Kb)

Hammer F., Flores H., Lilly S.J., Crampton D., Le Fevre O., Rola C., Mallen-Ornelas G., Schade D., Tresse L. Astrophys. J. 481, 49 (1997) Canada-France Redshift Survey. XIV. Spectral properties of field galaxies up to z=1.

J/ApJ/481/95/  table1.dat table2.dat  (39Kb)

Mattox J.R., Schachter J., Molnar L., Hartman R.C., Patnaik A.R. Astrophys. J. 481, 95 (1997) The identification of EGRET sources with flat-spectrum radio sources.

J/ApJ/483/745/  table3.dat  (11Kb)

Neilsen E.H., Tsvetanov Z.I., Ford H.C. Astrophys. J. 483, 745 (1997) The surface brightness fluctuations and globular cluster population of NGC 4478.

J/ApJ/483/826/  fig1a.dat fig1b.dat fig2.dat  (101Kb)

Phelps R.L. Astrophys. J. 483, 826 (1997) (From AAS CD-ROM series, Volume 9, 1997 Lee E. Brotzman [ADS] 24-Jul-97) Berkeley 17: the oldest open cluster?

J/ApJ/484/394/ ecoa656.dat ecob448.dat ecob777.dat ecoc594.dat jpaa545.dat jpaa558.dat jpaa573.dat jpaa587.dat jpaa604.dat jpaa620.dat jpaa638.dat jpaa656.dat jpaa677.dat jpab545.dat jpab558.dat jpab573.dat jpab587.dat jpab604.dat jpab620.dat jpab638.dat jpab656.dat jpab677.dat jpac545.dat jpac558.dat jpac573.dat jpac587.dat jpac604.dat jpac620.dat jpac638.dat jpac656.dat jpac677.dat photom.dat polar.dat vrad.dat (1.2 Mb) Elias N.M.II, Wilson R.E., Olson E.C., Aufdenberg J.P., Guinan E.F., Guedel M., Van Hamme W.V., Stevens H.L. Astrophys. J. 484, 394 (1997) (From AAS CD-ROM series, Volume 9, 1997 Lee E. Brotzman [ADS] 24-Jul-97) New Perspectives on AX Monocerotis

J/ApJ/488/224/  table3.dat table4.dat  (63Kb)

Kuchar T.A., Clark F.O. Astrophys. J. 488, 224 (1997) (From AAS CD-ROM series, Volume 9, 1997 Lee E. Brotzman [ADS] 29-Jul-97) All-Sky 4.85 GHz flux measurements of H II regions

J/ApJ/490/328/  table2.dat  (14Kb)

Burkholder V., Massey P., Morrell N.

Astrophys. J. 490, 328 (1997) (From AAS CD-ROM series, Volume 9, 1997 Lee E. Brotzman [ADS] 27-Jul-97) The ”mass discrepancy” for massive stars: tests of models using spectroscopic binaries

J/ApJ/491/L31/  table1.dat table1.tex  (9Kb)

Garzon F., Lopez-Corredoira M., Hammersley P., Mahoney T.J., Calbet X., Beckman J.E. Astrophys. J. 491, L31 (1997) A major star formation region in the receding tip of the stellar Galactic bar

J/ApJ/491/93/  grbcat.dat  (141Kb)

Larson S.B., McLean I.S. Astrophys. J. 491, 93 (1997) Extragalactic content of gamma-ray burst localizations

J/ApJ/492/62/  table1.dat table2.dat  (21Kb)

Ciardullo R., Jacoby G.H., Feldmeier J.J., Bartlett R.E. Astrophys. J. 492, 62 (1998) The planetary nebula luminosity function of M 87 and the intracluster stars of Virgo.

J/ApJ/496/39/  table1.dat table2.dat  (59Kb)

Zabludoff A.I., Mulchaey J.S. Astrophys. J. 496, 39 (1998) The properties of poor groups of galaxies. I. Spectroscopic survey and results.

J/ApJ/497/188/  table3.dat table4.dat  (71Kb)

Couch W.J., Barger A.J., Smail I., Ellis R.S., Sharples R.M. Astrophys. J. 497, 188 (1998) Morphological studies of the galaxy populations in distant “Butcher-Oemler” clusters with the Hubble Space Telescope. II. AC 103, AC 118, and AC 114 at z = 0.31.

J/ApJ/497/330/  table2.dat table3.dat table4.dat table5.dat  (26Kb)

Mannings V., Barlow M.J. Astrophys. J. 497, 330 (1998) Candidate main-sequence stars with debris disks: a new sample of Vega-like sources.

J/ApJ/497/736/  table1a.dat table5.dat table6.dat  (39Kb)

Herbig G.H. Astrophys. J. 497, 736-758 (1998) The young cluster IC 348.

J/ApJ/498/195/  table1.dat  (28Kb)

Fisher D., Fabricant D., Franx M., Van Dokkum P. Astrophys. J. 498, 195 (1998) A spectroscopic survey of the galaxy cluster CL 1358+62 at z = 0.328.

J/ApJ/499/112/  table2.dat  (35Kb)

Brinchmann J., Abraham R., Schade D., Tresse L., Ellis R.S., Lilly S., Le Fevre O., Glazebrook K., Hammer F., Colless M., Crampton D., Broadhurst T. Astrophys. J. 499, 112 (1998) Hubble space telescope imaging of the CFRS and LDSS redshift surveys. I. Morphological properties.

J/ApJ/499/517/  table1.dat table3.dat  (91Kb)

Pan J., Albert S., Sastry K.V.L.N., Herbst E., De Lucia F.C. Astrophys. J. 499, 517 (1998) The millimeter- and submillimeter-wave spectrum of ethylene oxide (c-C_{2}H_{4}O)

J/ApJ/500/554/  table3.dat table12.dat  (13Kb)

Odenwald S., Newmark J., Smoot G. Astrophys. J. 500, 554 (1998) A study of external galaxies detected by the COBE diffuse infrared background experiment.

J/ApJ/500/642 See IX/19/

J/ApJ/500/714/  table1.dat  (15Kb)

Van Dokkum P.G., Franx M., Kelson D.D., Illingworth G.D., Fisher D., Fabricant D. Astrophys. J. 500, 714 (1998) The color-magnitude relation in CL 1358+62 at z = 0.33: evidence for significant evolution in the S0 population.

J/ApJ/500/1059/  table1.dat table3.dat  (533Kb)

Groner P., Albert S., Herbst E., De Lucia F.C. Astrophys. J. 500, 1059 (1998) Dimethyl ether: laboratory assignments and predictions through 600 GHz.

J/ApJ/501/153/  table2.dat table3.dat  (106Kb)

Massey P. Astrophys. J. 501, 153 (1998) Evolved massive stars in the local group. I. Identification of red supergiants in NGC 6822, M31, and M33.

J/ApJ/502/558/  notes.dat table3.dat  (22Kb)

Vikhlinin A., McNamara B.R., Forman W., Jones C., Quintana H., Hornstrup A. Astrophys. J. 502, 558 (1998) A catalog of 200 galaxy clusters serendipitously detected in the ROSAT PSPC pointed observations.

J/ApJ/504/27/  table1.dat  (12Kb)

Markevitch M. Astrophys. J. 504, 27 (1998) The L_{X}-T Relation and Temperature Function for Nearby Clusters Revisited

J/ApJ/504/113/  table2.dat table3.dat  (33Kb)

Brown T.M., Ferguson H.C., Stanford S.A., Deharveng J. Astrophys. J. 504, 113 (1998) Color-luminosity relations for the resolved hot stellar populations in the centers of M31 and M32.

J/ApJ/505/50/  table2a.dat table2b.dat  (22Kb)

Bershady M.A., Lowenthal J.D., Koo D.C. Astrophys. J. 505, 50 (1998) Near-infrared galaxy counts to J and K 24 as a function of image size.

J/ApJ/505/793/  table2.dat table9.dat table10.dat table11.dat table12.dat  (56Kb)

Massey P., Johnson O. Astrophys. J. 505, 793 (1998) Evolved massive stars in the local group. II. A new survey for Wolf-Rayet stars in M33 and its implications for massive star evolution: evidence of the “Conti scenario” in action.

J/ApJ/507/585/  table1.dat table2.dat  (97Kb)

Marleau F.R., Simard L. Astrophys. J. 507, 585 (1998) Quantitative morphology of galaxies in the Hubble Deep Field.

J/ApJ/507/655/ note5.dat note12.dat table5.dat table6.dat table7.dat table8.dat table9.dat table12.dat table13.dat table14.dat table15.dat table16.dat table17.dat table18.dat table19.dat table20.dat (72Kb) Ferrarese L., Bresolin F., Kennicutt R.C.Jr, Saha A., Stetson P.B., Freedman W.L., Mould J.R., Madore B.F., Sakai S., Ford H.C., Gibson B.K., Graham J.A., Han M., Hoessel J.G., Huchra J., Hughes S.M., Illingworth G.D., Phelps R., Prosser C.F., Silbermann N.A. Astrophys. J. 507, 655 (1998) The Hubble Space Telescope Key Project on the extragalactic distance scale. XII. The discovery of cepheids and a new distance to NGC 2541.

J/ApJ/508/200/  table1.dat table2.dat table4.dat table5.dat  (30Kb)

Tripp T.M., Lu L., Savage B.D. Astrophys. J. 508, 200 (1998) The relationship between galaxies and low-redshift weak Lyαabsorbers in the directions of H1821+643 and PG 1116+215.

J/ApJ/508/314/  table4.dat  (10Kb)

Mukherjee S., Feigelson E.D., Babu J.G., Murtagh F., Fraley C., Raftery A. Astrophys. J. 508, 314 (1998) Three types of gamma-ray bursts.

J/ApJ/508/397 See III/240/

J/ApJ/508/491/  table2.dat table3.dat table5.dat  (44Kb)

Stetson P.B., Saha A., Ferrarese L., Rawson D.M., Ford H.C., Freedman W.L., Gibson B.K., Graham J.A., Harding P., Han M., Hill R.J., Hoessel J.G., Huchra J.P., Hughes S.M.G., Illingworth G.D., Kelson D.D., Kennicutt R.C.Jr, Madore B.F., Mould J.R., Phelps R.L., Sakai S., Silbermann N.A., Turner A. Astrophys. J. 508, 491 (1998) The extragalactic distance scale key project. XVI. Cepheid variables in an inner field of M101.

J/ApJ/509/309/  table3.dat  (12Kb)

Cabanac R.A., Borra E.F., Beauchemin M. Astrophys. J. 509, 309 (1998) A search for peculiar objects with the NASA orbital debris observatory 3 meter liquid mirror telescope.

J/ApJ/510/64/  table1.dat table2.dat  (11Kb)

Haarsma D. B., Hewitt J. N., Lehar J., Burke B. F. Astrophys. J. 510, 64 (1999) The Radio Wavelength Time Delay of Gravitational Lens 0957+561

J/ApJ/510/104/  table3.dat  (16Kb)

Bresolin F., Kennicutt R. C., Garnett D. R. Astrophys. J. 510, 104 (1999) The Ionizing Stars of Extragalactic H II Regions

J/ApJ/510/637/  table1.dat  (11Kb)

Hunt L. K., Malkan M. A., Moriondo G., Salvati M. Astrophys. J. 510, 637 (1999) The Disks of Galaxies with Seyfert and Starburst Nuclei. II. Near-Infrared Structural Properties

J/ApJ/510/659/  table4.dat  (17Kb)

Peterson B. M., Barth A. J., Berlind P., Bertram R., Bischoff K., Bochkarev N. G., Burenkov A. N., Cheng F.-Z., Dietrich M., Pilippenko A. V., Giannuzzo E., Ho L. C., Huchra J. P., Hunley J., Kaspi S., Kollatschny W., Leonard D. C., Malkov Y. F., Matheson T., Mignoli M., Nelson B., Papaderos P., Peters J., Pogge R. W., Pronik V. I., Sergeev S. G., Sergeeva E. A., Shapovalova A. I., Stirpe G. M., Tokarz S., Wagner R. M., Wanders I., Wei J.-Y., Wilkes B. J., Wu H., Xue S.-J., Zou Z.-L. Astrophys. J. 510, 659 (1999) Steps toward Determination of the Size and Structure of the Broad-Line Region in Active Galactic Nuclei. XV. Long-Term Optical Monitoring of NGC 5548

J/ApJ/511/65/  refs.dat table3.dat table4.dat  (78Kb)

Jones C., and Forman W. Astrophys. J. 511, 65 (1999) Einstein Observatory Images of Clusters of Galaxies

J/ApJ/511/595/  table1.dat table2.dat table3.dat  (18Kb)

Henriksen M., Cousineau S. Astrophys. J. 511, 595 (1999) An X-Ray Survey of Galaxies in Pairs

J/ApJ/511/612/  table1.dat  (13Kb)

Goldschmidt P., Kukula M. J., Miller L., Dunlop J. S. Astrophys. J. 511, 612 (1999) A Comparison of the Optical Properties of Radio-loud and Radio-quiet Quasars

J/ApJ/511/639/  table1.dat  (10Kb)

Izotov Y.I., Thuan T.X. Astrophys. J. 511, 639 (1999) Heavy-element abundances in blue compact galaxies

J/ApJ/512/48/  table4.dat table5.dat table6a.dat table6b.dat  (36Kb)

Gibson B.K., Hughes S.M.G., Stetson P.B., Freedman W.L., Kennicutt R.C.Jr, Mould J.R., Bresolin F., Ferrarese L., Ford H.C., Graham J.A., Han M., Harding P., Hoessel J.G., Huchra J.P., Illingworth G.D., Kelson D.D., Macri L.M., Madore B. F., Phelps R.L., Prosser C.F., Saha A., Sakai S., Sebo K.M., Silbermann N.A., Turner A.M. Astrophys. J. 512, 48 (1999) The Hubble Space Telescope key project on the extragalactic distance scale. XVII. The Cepheid distance to NGC 4725

J/ApJ/512/125/  table1.dat  (16Kb)

Schmidt G.D., Hines D.C. Astrophys. J. 512, 125 (1999) The polarization of broad absorption line QSOs

J/ApJ/513/34/  table4.dat table5.dat  (199Kb)

Fernandez-Soto A., Lanzetta K.M., Yahil A. Astrophys. J. 513, 34 (1999) A new catalog of photometric redshifts in the Hubble Deep Field.

J/ApJ/513/168/  table2.dat  (21Kb)

Garnett D.R., Shields G.A., Peimbert M., Torres-Peimbert S., Skillman E.D., Dufour R.J., Terlevich E., Terlevich R.J. Astrophys. J. 513, 168 (1999) Carbon in spiral galaxies from Hubble space telescope spectroscopy

J/ApJ/513/733/  table2.dat  (9Kb)

Kundu A., Whitmore B.C., Sparks W.B., Macchetto F.D., Zepf S.E., Ashman K.M.

Astrophys. J. 513, 733 (1999) The globular cluster system in the inner region of M87.

J/ApJ/514/148/  notes.dat refs.dat table1.dat  (34Kb)

De Grandi S., Bohringer H., Guzzo L., Molendi S., Chincarini G., Collins C., Cruddace R., Neumann D., Schindler S., Schuecker P., Voges W. Astrophys. J. 514, 148 (1999) A flux-limited sample of bright clusters of galaxies from the southern part of the ROSAT all-sky survey: the catalog and log n-log s

J/ApJ/514/614/  table5.dat table6.dat tablea1.dat tablea2.dat  (104Kb)

Kelson D.D., Illingworth G.D., Saha A., Graham J.A., Stetson P.B., Freedman W.L., Kennicutt R.C., Mould J.R., Ferrarese L., Huchra J.P., Madore B.F., Prosser C.F. Bresolin F., Ford H.C., Gibson B. K., Hoessel J.G., Hughes S.M.G., Macri L.M., Sakai S., Silbermann M.A. Astrophys. J. 514, 614 (1999) The Hubble Space Telescope Key Project on the Extragalactic Distance Scale. XIX. The Discovery of Cepheids in and a New Distance to NGC 3198

J/ApJ/515/610/  table8.dat table9.dat  (20Kb)

Beaulieu S.F., Dopita M.A., Freeman K.C. Astrophys. J. 515, 610 (1999) A survey of planetary nebulae in the southern Galactic Bulge.

J/ApJ/516/693/  table1.dat  (16Kb)

Honma M. Astrophys. J. 516, 693 (1999) The mass-to-light ratio of binary galaxies

J/ApJ/517/148/  table1.dat  (17Kb)

Flores H., Hammer F., Thuan T.X., Cesarsky C., Desert F.X., Omont A., Lilly S .J., Eales S., Crampton D., Le Fevre O. Astrophys. J. 517, 148 (1999) 15 micron Infrared Space Observatory observations of the 1415+52 Canada-France redshift survey field: the cosmic star formation rate as derived from deep ultraviolet, optical, mid-infrared, and radio photometry

J/ApJ/517/264/  table1.dat  (13Kb)

Arce H.G., Goodman A.A. Astrophys. J. 517, 264 (1999) An extinction study of the Taurus dark cloud complex

J/ApJ/517/282/  table1.dat  (16Kb)

Huang M., Bania T.M., Bolatto A., Chamberlin R.A., Ingalls J.G., Jackson J.M., Lane A.P., Stark A.A., Wilson R.W., Wright G.A. Astrophys. J. 517, 282 (1999) Atomic carbon observations of southern hemisphere H II region

J/ApJ/517/650/  table2.dat  (15Kb)

Ryden B.S., Terndrup D.M., Pogge R.W., Lauer T.R. Astrophys. J. 517, 650 (1999) Detailed surface photometry of dwarf elliptical and dwarf S0 galaxies in the Virgo cluster

J/ApJ/518/69/  table1.dat table2.dat  (43Kb)

Mahdavi A., Geller M.J., Bohringer H., Kurtz M.J., Ramella M. Astrophys. J. 518, 69 (1999) The dynamics of poor systems of galaxies.

J/ApJ/518/262/  table1.dat  (1.7 Mb)

Sandquist E.L., Bolte M., Langer G.E., Hesser J.E., De Oliveira C.M. Astrophys. J., 518, 262-283 (1999) Wide-field CCD photometry of the globular cluster M30.

J/ApJ/518/656/  table3.dat  (17Kb)

Ueda Y., Takahashi T., Inoue H., Tsuru T., Sakano M., Ishisaki Y., Ogasaka Y., Makishima K., Yamada T., Akiyama M., Ohta K. Astrophys. J. 518, 656 (1999) logN-logS relations and spectral properties of sources from the ASCA Large Sky Survey: their implications for the origin of the cosmic X-ray background (CXB)

J/ApJ/518/859/  table1.dat  (15Kb)

Keenan P.C., Barnbaum C. Astrophys. J. 518, 859 (1999)

Revision and calibration of MK luminosity classes for cool giants by HIPPARCOS parallaxes

J/ApJ/519/1/  table1.dat table2.dat  (13Kb)

Steidel C.C., Adelberger K.L., Giavalisco M., Dickinson M., Pettini M. Astrophys. J. 519, 1 (1999) Lyman-break galaxies at z 4 and the evolution of the ultraviolet luminosity density at high redshift

J/ApJ/519/533/  table1.dat  (23Kb)

David L.P., Forman W., Jones C. Astrophys. J. 519, 533 (1999) ROSAT PSPC observations of the richest (R =2) ACO clusters

J/ApJ/519/697/  table2.dat  (10Kb)

Osamura Y., Fukuzawa K., Terzieva R., Herbst E. Astrophys. J. 519, 697 (1999) A molecular orbital study of the HC_{3}NH^{+}+e^{−} dissociative recombination and its role in the production of cyanoacetylene isomers in interstellar clouds

J/ApJ/521/577/  table1.dat  (50Kb)

Hodge P.W., Dolphin A.E., Smith T.R. Mateo M. Astrophys. J. 521, 577 (1999) Hubble space Telescope Studies of the WLM galaxy. I. The age and metallicity of the globular cluster

J/ApJ/521/682/  table2.dat  (65Kb)

Abt H.A., Willmarth D.W. Astrophys. J. 521, 682 (1999) Binaries in the Praesepe and Coma star clusters and their implications for binary evolution

J/ApJ/522/113/  table2.dat table3.dat  (24Kb)

Veilleux S., Kim D.-C., Sanders D.B. Astrophys. J. 522, 113 (1999) Optical spectroscopy of the IRAS 1Jy sample of ultraluminous infrared galaxies.

J/ApJ/522/802/  table2.dat table3.dat table4.dat  (63Kb)

Saha A., Sandage A., Tammann G.A., Labhardt L., Macchetto F.D., Panagia N. Astrophys. J. 522, 802 (1999) Cepheid calibration of the peak brightness of type Ia supernovae. IX. SN1989B in NGC 3627

J/ApJ/522/846/  table2.dat  (72Kb)

Xie G.Z., Li K.H., Zhang X., Bai J.M., Liu W.W. Astrophys. J. 522, 846 (1999) Optical monitoring sample of the GeV gamma-ray-loud blazars

J/ApJ/523/248/  table1.dat table2.dat  (19Kb)

Ott T., Eckart A., Genzel R. Astrophys. J. 523, 248 (1999) Variable and embedded stars in the Galactic center.

J/ApJ/523/540/  table4.dat table5.dat table6.dat table7.dat  (46Kb)

Sakai S., Ferrarese L., Kennicutt R.C., Graham J.A., Silbermann N.A., Mould J .R., Freedman W.L., Bresolin F., Ford H.C., Gibson B.K., Han M., Harding P., Hoessel J.G., Huchra J.P., Hughes S.M., Illingworth G.D., Kelson D., Macri L., Madore B.F., Phelps R.L., Saha A., Sebo K.M., Stetson P.B., Turner A. Astrophys. J. 523, 540 (1999) The Hubble SPACE TELESCOPE extragalactic distance scale key project. XXIII. The discovery of cepheids in NGC 3319.

J/ApJ/524/536/  table2.dat  (53Kb)

Impey C.D., Petry C.E., Flint K.P. Astrophys. J. 524, 536 (1999) A study of Lyαquasar absorbers in the nearby universe

J/ApJ/525/127/  refs.dat table2.dat  (24Kb)

Laurent-Muehleisen S.A., Kollgaard R.I., Feigelson E.D., Brinkmann W., Siebert J. Astrophys. J. 525, 127 (1999) The RGB sample of intermediate BL Lacertae objects

J/ApJ/525/420/  table1.dat table2.dat  (13Kb)

Gies D.R., Kambe E., Josephs T.S., Bagnuolo W.G., Choi Y.J., Gudehus D., Guyton K.M., Hartkopf W.I., Hildebrand J.L., Kaye A.B., Mason B.D., Riddle R.L., Sowers J.W., Turner N.H., Wilson J.W., Xiong Y. Astrophys. J. 525, 420 (1999) Ultraviolet and optical line profile variations in the spectrum of ϵPersei

J/ApJ/525/466/  table1.dat  (16Kb)

Luhman K.L. Astrophys. J. 525, 466 (1999) Young low-mass stars and brown dwarfs in IC 348

J/ApJ/526/544/  table2.dat  (24Kb)

Tayal S.S., Gupta G.P. Astrophys. J. 526, 544 (1999) Collision strengths for electron collision excitation of fine-structure levels in S III

J/ApJ/526/788/  table1.dat table2.dat  (27Kb)

Lee C.W., Myers P.C., Tafalla M. Astrophys. J. 526, 788 (1999) A survey of infall motions toward starless cores. I. CS (2-1) and N_{2}H^{+} (1-0) observations

J/ApJ/527/199/  table2a.dat table2b.dat table2c.dat  (2.5 Mb)

Johnson J.A., Bolte M., Stetson P.B., Hesser J.E., Somerville R.S. Astrophys. J. 527, 199 (1999) Hubble space telescope observations of the oldest star clusters in the large Magellanic cloud.

J/ApJ/527/573/  refs.dat table1.dat table2.dat table4.dat  (39Kb)

Kobayashi, C., Arimoto N. Astrophys. J. 527, 573 (1999) Gradients of Absorption-Line Strengths in Elliptical Galaxies

J/ApJ/527/879/  table1.dat  (28Kb)

Allende Prieto C., Garcia Lopez R.J., Lambert D.L., Gustafsson B. Astrophys. J. 527, 879 (1999) A consistency test of spectroscopic gravities for late-type stars

J/ApJ/528/123/  table1.dat  (12Kb)

Tanaka I., Yamada T., Aragon-Salamanca A., Kodama T., Miyaji T., Ohta K., Arimoto N. Astrophys. J. 528, 123 (2000) A rich cluster of galaxies near the quasar B2 1335+28 at z=1.1: color distribution and star formation properties.

J/ApJ/529/723/  appen.dat  (22Kb)

Gibson B.K., Stetson P.B., Freedman W.L., Mould J.R., Kennicutt R.C.Jr, Huchra J.P., Sakai S., Graham J.A., Fassett C.I., Kelson D.D., Ferrarese L., Hughes S.M.G., Illingworth G.D., Macri L.M., Madore B.F., Sebo K.M., Silbermann N.A. Astrophys. J. 529, 723 (2000) The Hubble space telescope key project on the extragalactic distance scale. XXV. A recalibration of Cepheid distances to type Ia supernovae and the value of the Hubble constant.

J/ApJ/530/783/  table2.dat  (23Kb)

Westin J., Sneden C., Gustafsson B., Cowan J.J. Astrophys. J. 530, 783 (2000) The r-Process-enriched Low-Metallicity Giant HD 115444

J/ApJ/532/308/  table1.dat  (521Kb)

Brown T.M., Bowers C.W., Kimble R.A., Sweigart A.V., Ferguson H.C. Astrophys. J. 532, 308 (2000) Detection and Photometry of Hot Horizontal Branch Stars in the Core of M32

J/ApJ/533/611/  table2.dat table5.dat  (34Kb)

Richards E.A. Astrophys. J. 533, 611 (2000) The nature of radio emission from distant galaxies: the 1.4 GHz observations.

J/ApJ/534/114/  table2.dat table3.dat table4.dat  (35Kb)

Mahdavi A., Bohringer H., Geller M.J., Ramella M. Astrophys. J. 534, 114 (2000) The RASSCALS: an X-ray and optical study of 260 galaxy groups.

J/ApJ/535/325 See III/240/

J/ApJ/538/29/  table2a.dat table2b.dat table3.dat  (47Kb)

Cohen J.G., Hogg D.W., Blandford R., Cowie L.L., Hu E., Songaila A., Shopbell P., Richberg K. Astrophys. J. 538, 29 (2000) Caltech faint galaxy redshift survey. X. A redshift survey in the region of the Hubble Deep Field North.

J/ApJ/538/493/ errorwf.dat niphot.dat niphotoz.dat table2.dat table3.dat table6.dat table8.dat wfphot.dat wfphotoz.dat (561Kb)

Yahata N., Lanzetta K.M., Chen H.-W., Fernandez-Soto A., Pascarelle S.M., Yahil A., Puetter R.C. Astrophys. J. 538, 493 (2000) Photometry and photometric redshifts of faint galaxies in the Hubble Deep Field South NICMOS field.

J/ApJ/539/136/  table1.dat table2.dat  (57Kb)

Zabludoff A.I., Mulchaey J.S. Astrophys. J. 539, 136 (2000) The properties of poor groups of galaxies. III. The galaxy luminosity function.

J/ApJ/540/236/  table1.dat  (62Kb)

Hillenbrand L.A., Carpenter J.M. Astrophys. J. 540, 236 (2000) Constraints on the stellar/substellar mass function in the inner Orion Nebula Cluster.

J/ApJ/540/1016/  table1.dat  (67Kb)

Luhman K.L., Rieke G.H., Young E.T., Cotera A.S., Chen H., Rieke M.J., Schneider G., Thompson R.I. Astrophys. J. 540, 1016 (2000) The initial mass function of low-mass stars and brown dwarfs in young clusters.

J/ApJ/541/841/  table3a.dat table3b.dat  (219Kb)

Abt H.A., Corbally C.J. Astrophys. J. 541, 841 (2000) The maximum age of Trapezium systems.

J/ApJ/541/977/  table1.dat table4.dat table6.dat  (40Kb)

Najita J.R., Tiede G.P., Carr J.S. Astrophys. J. 541, 977 (2000) From stars to superplanets: the low-mass initial mass function in the young cluster IC 348.

J/ApJ/542/257/  table1.dat table2.dat table3.dat table4.dat table5.dat  (35Kb)

Alcock C., Allsman R., Alves D.R., Axelrod T., Becker A., Bennett D., Clement C., Cook K.H., Drake A., Freeman K., Geha M., Griest K., Kovacs G., Kurtz D.W., Lehner M., Marshall S., Minniti D., Nelson C., Peterson B., Popowski P., Pratt M., Quinn P., Rodgers A., Rowe J., Stubbs C., Sutherland W., Tomaney A., Vandehei T., Welch D.L. Astrophys. J. 542, 257 (2000) The MACHO project Large Magellanic Cloud variable-star inventory. IX. Frequency analysis of the firstovertone RR Lyrae stars and the indication for nonradial pulsations.

J/ApJ/542/673/  table2.dat table3.dat table5.dat  (31Kb)

Fasano G., Poggianti B.M., Couch W.J., Bettoni D., Kjaergaard P., Moles M. Astrophys. J. 542, 673 (2000) The evolution of the galactic morphological types in clusters.

J/ApJ/543/178/  table2.dat table3.dat table4.dat table5.dat table6.dat table7.dat  (1.6 Mb)

Giuricin G., Marinoni C., Ceriani L., Pisani A. Astrophys. J. 543, 178 (2000) Nearby optical galaxies: selection of the sample and identification of groups.

J/ApJ/543/552/  refs.dat table1.dat table2.dat table3.dat table4.dat table5.dat table6.dat  (66Kb)

Storrie-Lombardi L.J., Wolfe A.M. Astrophys. J. 543, 552 (2000) Surveys for z 3 damped Lyαabsorption systems: the evolution of neutral gas.

J/ApJ/543/1016/  notes1.dat table1.dat table2.dat table3.dat  (21Kb)

Brosius J.W., Thomas R.J., Davila J.M., Landi E. Astrophys. J. 543, 1016 (2000) Analysis of a solar active region extreme-ultraviolet spectrum from SERTS-97.

J/ApJ/546/694/  table1.dat  (35Kb)

Thompson R.I., Weymann R.J., Storrie-Lombardi L.J. Astrophys. J. 546, 694 (2001) Star formation history in the NICMOS Northern Hubble Deep Field.

J/ApJ/546/1006/  table2.dat  (160Kb)

Barrado y Navascues D., Stauffer J.R., Bouvier J., Martin E.L. Astrophys. J. 546, 1006 (2001) From the top to the bottom of the main sequence: a complete mass function of the young open cluster M35.

J/ApJ/548/127/  catalog.dat  (17.1 Mb)

Teplitz H.I., Hill R.S., Malumuth E.M., Collins N.R., Gardner J.P., Palunas P., Woodgate B.E. Astrophys. J. 548, 127 (2001)

Wide-field imaging of the Hubble Deep Field-South region. II. The evolution of galaxy clustering at z 1.

J/ApJ/549/172/  table2.dat table3.dat  (23Kb)

Tran K.-V.H., Simard L., Zabludoff A.I., Mulchaey J.S. Astrophys. J. 549, 172 (2001) The galaxy populations of X-ray-detected, poor groups.

J/ApJ/549/441/  table1.dat  (12Kb)

Schulz N.S., Canizares C., Huenemoerder D., Kastner J.H., Taylor S.C., Bergstroem E.J. Astrophys. J. 549, 441 (2001) Chandra observations of variable embedded X-ray sources in Orion. I. Resolving the Orion Trapezium.

J/ApJ/549/578/  table1.dat table2.dat table3.dat table4.dat  (48Kb)

Degioia-Eastwood K., Throop H., Walker G., Cudworth K.M. Astrophys. J. 549, 578 (2001) The star formation history of Trumpler 14 and Trumpler 16.

J/ApJ/549/820/  table1.dat  (59Kb)

Couch W.J., Balogh M.L., Bower R.G., Smail I., Glazebrook K., Taylor M. Astrophys. J. 549, 820 (2001) A low global star formation rate in the rich galaxy cluster AC 114 at z=0.32.

J/ApJ/551/111/  table1.dat  (36Kb)

Ellis R.S., Abraham R.G., Dickinson M. Astrophys. J. 551, 111 (2001) The relative star formation histories of spiral bulges and elliptical galaxies in the Hubble Deep Fields.

J/ApJ/551/852/  table1.dat  (1.3 Mb)

Heyer M.H., Carpenter J.M., Snell R.L. Astrophys. J. 551, 852 (2001) The equilibrium state of molecular regions in the outer Galaxy.

J/ApJ/551/973/  table2.dat table3.dat table4.dat table5.dat  (61Kb)

Saha A., Sandage A., Thim F., Labhardt L., Tammann G.A., Christensen J., Panagia N., Macchetto F.D. Astrophys. J. 551, 973 (2001) Cepheid calibration of the peak brightness of type Ia supernovae. X. SN 1991T in NGC 4527.

J/ApJ/552/289/  table1.dat table2.dat table3.dat  (47Kb)

Alard C., Blommaert J.A.D.L., Cesarsky C., Epchtein N., Felli M., Fouque P., Ganesh S., Genzel R., Gilmore G., Glass I.S., Habing H., Omont A., Perault M., Price S., Robin A., Schultheis M., Simon G., Van Loon J.T., Alcock C., Allsman R.A., Alves D.R., Axelrod T.S., Becker A.C., Bennett D.P., Cook K.H., Drake A.J., Freeman K.C., Geha M., Griest K., Lehner M.J., Marshall S.L., Minniti D., Nelson C., Peterson B.A., Popowski P., Pratt M.R., Quinn P.J., Sutherland W., Tomaney A.B., Vandehei T., Welch D.L. Astrophys. J. 552, 289 (2001) Mass-losing semiregular variable stars in Baade’s windows.

J/ApJ/554/L129/  galaxy.dat system.dat  (32Kb)

Mahdavi A., Geller M.J. Astrophys. J. 554, L129 (2001) The L_{X}-σrelation for galaxies and clusters of galaxies.

J/ApJ/554/202/  table3.dat table4.dat table5.dat table6.dat  (21Kb)

Immler S., Wang Q.D. Astrophys. J. 554, 202 (2001) ROSAT X-ray observations of the spiral galaxy M81.

J/ApJ/554/742/  table4.dat table6.dat table8.dat table9.dat table10.dat  (42Kb)

Hornschemeier A.E., Brandt W.N., Garmire G.P., Schneider D.P., Barger A.J., Broos P.S., Cowie L.L., Townsley L.K., Bautz M.W., Burrows D.N., Chartas G., Feigelson E.D., Griffiths R.E., Lumb D., Nousek J.A., Ramsey L.W., Sargent W.L.W. Astrophys. J. 554, 742 (2001) The Chandra Deep Survey of the Hubble Deep Field-North area. II. Results from the Caltech Faint Field Galaxy Redshift Survey area.

J/ApJ/554/803/  catalog.dat  (190Kb)

Yun Min S., Reddy Naveen A., Condon J.J. Astrophys. J. 554, 803 (2001) Radio Properties of Infrared-selected Galaxies in the IRAS 2 Jy Sample

J/ApJ/555/932/  notes.dat table2.dat table3.dat table4.dat table5.dat table6.dat table7.dat  (35Kb)

Spangler C., Sargent A.I., Silverstone M.D., Becklin E.E., Zuckerman B.

Astrophys. J. 555, 932 (2001) Dusty debris around solar-type stars: temporal disk evolution.

J/ApJ/559/307/  table1.dat  (11Kb)

Johnstone D., Fich M., Mitchell G.F., Moriarty-Schieven G. Astrophys. J. 559, 307 (2001) Large area mapping at 850 microns. III. Analysis of the clump distribution in the Orion B molecular cloud.

J/ApJ/559/606/  table3.dat  (6Kb)

Carter B.J., Fabricant D.G., Geller M.J., Kurtz M.J., McLean B. Astrophys. J. 559, 606 (2001) Star formation in a complete spectroscopic survey of galaxies.

J/ApJ/559/654/  object.dat table3.dat  (25Kb)

Chen H.-W., Lanzetta K.M., Webb J.K., Barcons X. Astrophys. J. 559, 654 (2001) The gaseous extent of galaxies and the origin of Lyαabsorption systems. V. Optical and near-infrared photometry of lyα-absorbing galaxies at z 1.

J/ApJ/559/812/  table1.dat  (35Kb)

Hanes D.A., Cote P., Bridges T.J., McLaughlin D.E., Geisler D., Harris G.L.H., Hesser J.E., Lee M.G. Astrophys. J., 559, 812-827 (2001) Dynamics of the globular cluster system associated with M87 (NGC 4486). I. New CFHT MOS spectroscopy and the composite database.

J/ApJ/559/851/  table1.dat table4.dat  (18Kb)

Williams B.F., Hodge P.W. Astrophys. J. 559, 851 (2001) Identification and analysis of young star cluster candidates in M31.

J/ApJ/559/864/  table1.dat table2.dat table3.dat table4.dat table5.dat table6.dat  (18Kb)

Johnson K.E., Kobulnicky H.A., Massey P., Conti P.S. Astrophys. J. 559, 864 (2001) A sample of clusters of extragalactic ultracompact H II regions.

J/ApJ/559/1060/  tables.dat  (18Kb)

Albrow M.D., Gilliland R.L., Brown T.M., Edmonds P.D., Guhathakurta P., Sarajedini A. Astrophys. J. 559, 1060 (2001) The frequency of binary stars in the core of 47 Tucanae.

J/ApJ/560/L131 See VII/228/

J/ApJ/560/566/  refs.dat table1.dat  (382Kb)

Kochanek C.S., Pahre M.A., Falco E.E., Huchra J.P., Mader J., Jarrett T.H., Chester T., Cutri R., Schneider S.E. Astrophys. J. 560, 566 (2001) The K-band galaxy luminosity function.

J/ApJ/560/675/  table2.dat  (20Kb)

Kraft R.P., Kregenow J.M., Forman W.R., Jones C., Murray S.S. Astrophys. J. 560, 675 (2001) Chandra observations of the X-ray point source population in Centaurus A.

J/ApJ/561/106/  hdfnpz.dat hdfspz.dat refs.dat table1.dat  (179Kb)

Mortsell E., Gunnarsson C., Goobar A. Astrophys. J. 561, 106 (2001) Gravitational lensing of the farthest known supernova SN 1997ff.

J/ApJ/561/189/  table1.dat table2.dat table4.dat  (21Kb)

Pence W.D., Snowden S.L., Mukai K., Kuntz K.D. Astrophys. J. 561, 189 (2001) Chandra X-ray sources in M101.

J/ApJ/561/727/  table1.dat  (8Kb)

Zhang Q., Fall S.M., Whitmore B.C. Astrophys. J. 561, 727 (2001) A multiwavelength study of the young star clusters and interstellar medium in the Antennae galaxies.

J/ApJ/562/254/  table3.dat table4.dat table5.dat  (18Kb)

Donnelly R.H., Forman W., Jones C., Quintana H., Ramirez A., Churazov E., Gilfanov M. Astrophys. J. 562, 254 (2001) Merging binary clusters.

J/ApJ/562/303/  field1.dat field2.dat field3.dat field4.dat ground.dat ngc121.dat  (1.5 Mb)

Dolphin A.E., Walker A.R., Hodge P.W., Mateo M., Olszewski E.W., Schommer R.A., Suntzeff N.B. Astrophys. J. 562, 303 (2001) Old Stellar Populations of the Small Magellanic Cloud

J/ApJ/562/337/  table1.dat table2.dat table3.dat  (24Kb)

Alcock C., Allsman R.A., Alves D.R., Axelrod T.S., Becker A.C., Bennett D.P., Cook K.H., Drake A.J., Freeman K.C., Geha M., Griest K., Lehner M.J., Marshall S.L., Minniti D., Nelson C.A., Peterson B.A., Popowski P., Pratt M.R., Quinn P.J., Stubbs C.W., Sutherland W., Vandehei T., Welch D.L. Astrophys. J. 562, 337 (2001) Astrometry with the MACHO data archive. I. High proper motion stars toward the Galactic Bulge and Magellanic clouds.

J/ApJ/562/368/  table2.dat  (19Kb)

Brown T.M., Sweigart A.V., Lanz T., Landsman W.B., Hubeny I. Astrophys. J. 562, 368 (2001) Flash mixing on the white dwarf cooling curve: understanding hot horizontal branch anomalies in NGC 2808.

J/ApJ/562/446/  table2.dat table13.dat  (13Kb)

Rho J., Corcoran M.F., Chu Y.-H., Reach W.T. Astrophys. J. 562, 446 (2001) X-rays and protostars in the Trifid nebula.

J/ApJ/562/528/  table2.dat table3.dat  (23Kb)

Snider S., Allende Prieto C., von Hippel T., Beers T.C., Sneden C., Qu Y., Rossi S. Astrophys. J. 562, 528 (2001) Three-dimensional spectral classification of low-metallicity stars using artificial neural networks.

J/ApJ/563/L49/  table1.dat table2.dat  (10Kb)

Hulleman F., Tennant A.F., van Kerkwijk M.H., Kulkarni S.R., Kouveliotou C., Patel S.K. Astrophys. J. 563, L49 (2001) A possible faint near-infrared counterpart to the anomalous X-ray pulsar 1E 2259+586.

J/ApJ/563/80/  table2.dat  (213Kb)

Stern B.E., Tikhomirova Y., Kompaneets D., Svensson R., Poutanen J. Astrophys. J. 563, 80 (2001) An off-line scan of the BATSE daily records and a large uniform sample of gamma-ray bursts.

J/ApJ/563/629/  table1.dat table2.dat  (18Kb)

Nelson A.E., Gonzalez A.H., Zaritsky D., Dalcanton J.J. Astrophys. J. 563, 629 (2001) Cluster galaxy evolution from a new sample of galaxy clusters at 0.3 z 0.9.

J/ApJ/564/421/ table3.dat table4.dat table5.dat table6.dat table7.dat table10.dat table11.dat table12.dat table13.dat table15.dat table16.dat (91Kb) Burgasser A.J., Kirkpatrick J.D., Brown M.E., Reid I.N., Burrows A., Liebert J., Matthews K., Gizis J.E., Dahn C.C., Monet D.G., Cutri R.M., Skrutskie M.F. Astrophys. J. 564, 421 (2002) The spectra of T dwarfs. I. Near-infrared data and spectral classification.

J/ApJ/565/208/  table1.dat  (8Kb)

Patton D.R., Pritchet C.J., Carlberg R.G., Marzke R.O., Yee H.K.C., Hall P.B., Lin H., Morris S.L., Sawicki M., Shepherd C.W., Wirth G.D. Astrophys. J. 565, 208 (2002) Dynamically close galaxy pairs and merger rate evolution in the CNOC2 Redshift Survey.

J/ApJ/565/511/  notes.dat table1.dat table2.dat table3.dat  (63Kb)

Hoard D.W., Wachter S., Clark L.L., Bowers T.P. Astrophys. J. 565, 511 (2002) Infrared properties of cataclysmic variables in the 2 micron All-Sky Survey Second Incremental Data Release.

J/ApJ/566/103/  table1.dat table2.dat  (14Kb)

Nelson A.E., Gonzalez A.H., Zaritsky D., Dalcanton J.J. Astrophys. J. 566, 103 (2002) Revisiting brightest cluster galaxy evolution with the Las Campanas Distant Cluster Survey.

J/ApJ/566/744/  table3.dat table7.dat table9.dat  (15Kb)

Lewis A.D., Stocke J.T., Ellingson E., Gaidos E.J. Astrophys. J. 566, 744 (2002) New X-ray clusters in the EINSTEIN Extended Medium-Sensitivity Survey. I. Modifications to the X-ray luminosity function.

J/ApJ/566/880/  table2.dat table56.dat  (31Kb)

Giveon U., Sternberg A., Lutz D., Feuchtgruber H., Pauldrach A.W.A. Astrophys. J. 566, 880 (2002) The excitation and metallicity of galactic H II regions from Infrared Space Observatory SWS observations of mid-infrared fine-structure lines.

J/ApJ/566/974/  table1.dat table2.dat table3.dat  (42Kb)

Tsujimoto M., Koyama K., Tsuboi Y., Goto M., Kobayashi N. Astrophys. J. 566, 974 (2002) X-ray properties of young stellar objects in OMC-2 and OMC-3 from the Chandra X-ray observatory.

J/ApJ/566/993/  table1.dat table2.dat table3.dat  (23Kb)

Allen L.E., Myers P.C., di Francesco J., Mathieu R., Chen H., Young E. Astrophys. J. 566, 993 (2002) Hubble space Telescope/NICMOS imaging survey of the Ophiuchus (Lynds 1688) cluster.

J/ApJ/567/423/  table1.dat  (24Kb)

Kohno M., Koyama K., Hamaguchi K. Astrophys. J. 567, 423 (2002) Chandra observations of high-mass young stellar objects in the Monoceros R2 molecular cloud.

J/ApJ/569/23/  table1.dat table2.dat table3.dat  (22Kb)

Smith P.S., Schmidt G.D., Hines D.C., Cutri R.M., Nelson B.O. Astrophys. J. 569, 23 (2002) The optical polarization of near-infrared-selected quasi-stellar objects.

J/ApJ/569/689/  table1.dat table2.dat table3.dat table4.dat table6.dat table7.dat  (64Kb)

Donahue M., Scharf C.A., Mack J., Lee Y.P., Postman M., Rosati P., Dickinson M., Voit G.M., Stocke J.T. Astrophys. J. 569, 689 (2002) Distant cluster hunting. II. A comparison of X-ray and optical cluster detection techniques and catalogs from the ROSAT optical X-ray survey.

J/ApJ/569/720/  table2.dat table3.dat table4.dat table5.dat  (35Kb)

Girardi M., Manzato P., Mezzetti M., Giuricin G., Limboz F. Astrophys. J. 569, 720 (2002) Observational mass-to-light ratio of galaxy systems from poor groups to rich clusters.

J/ApJ/570/54 See VII/228/

J/ApJ/570/100/  table3.dat table4.dat table5.dat  (22Kb)

La Franca F., Fiore F., Vignali C., Antonelli A., Comastri A., Giommi P., Matt G., Molendi S., Perola G.C., Pompilio F. Astrophys. J. 570, 100 (2002) The BeppoSAX High-Energy Large-Area Survey. V. The nature of the hard X-ray source population and its evolution.

J/ApJ/571/56/  table2.dat  (8Kb)

Bernstein R.A., Freedman W.L., Madore B.F. Astrophys. J. 571, 56 (2002) The first detections of the extragalactic background light at 3000, 5500, and 8000Å. I. Results.

J/ApJ/571/136/  table2.dat  (15Kb)

Im M., Simard L., Faber S.M., Koo D.C., Gebhardt K., Willmer C.N.A., Phillips A., Illingworth G., Vogt N.P., Sarajedini V.L. Astrophys. J. 571, 136 (2002) The DEEP Groth Strip survey. X. Number density and luminosity function of field E/S0 galaxies at z 1.

J/ApJ/571/512/  notes.dat table1.dat  (15Kb)

Holberg J.B., Oswalt T.D., Sion E.M. Astrophys. J. 571, 512 (2002) A determination of the local density of white dwarf stars.

J/ApJ/571/665/  refs.dat spectra/* table1.dat table3.dat  (19.1 Mb)

Scott J., Bechtold J., Morita M., Dobrzycki A., Kulkarni V.P. Astrophys. J. 571, 665 (2002) A uniform analysis of the Lyαforest at z=0-5. V. The extragalactic ionizing background at low redshift.

J/ApJ/573/306/  notes.dat table1.dat table2.dat table3.dat table4.dat table5.dat  (41Kb)

Eck C.R., Cowan J.J., Branch D. Astrophys. J. 573, 306 (2002) A search for radio emission from supernovae with ages from about 1 week to more than 80 years.

J/ApJ/573/359/  table1.dat  (78Kb)

Abt H.A., Levato H., Grosso M. Astrophys. J. 573, 359 (2002) Rotational velocities of B stars.

J/ApJ/573/366/  table6.dat  (153Kb)

Muench A.A., Lada E.A., Lada C.J., Alves J. Astrophys. J. 573, 366 (2002) The luminosity and mass function of the Trapezium cluster: from B stars to the deuterium-burning limit.

J/ApJ/574/258/ notes2.dat notes3.dat table2.dat table3.dat table4.dat table5.dat table6.dat (348Kb) Feigelson E.D., Broos P., Gaffney III J.A., Garmire G., Hillenbrand L.A., Pravdo S.H., Townsley L., Tsuboi Y. Astrophys. J. 574, 258 (2002) X-ray-emitting young stars in the Orion nebula.

J/ApJ/575/354/  notes.dat table1.dat table2.dat  (40Kb)

Getman K.V., Feigelson E.D., Townsley L., Bally J., Lada C.J., Reipurth B. Astrophys. J. 575, 354 (2002) Chandra study of young stellar objects in the NGC 1333 star-forming cloud.

J/ApJ/576/880/  table1.dat table2.dat table3.dat  (361Kb)

Slesnick C.L., Hillenbrand L.A., Massey P. Astrophys. J. 576, 880 (2002) The star formation history and mass function of the double cluster h and χPersei.

J/ApJ/577/31/ table2.dat table3.dat table4.dat table5.dat table6.dat table7.dat table8.dat table9.dat table10.dat (41Kb) Ciardullo R., Feldmeier J.J., Jacoby G.H., Kuzio de Naray R., Laychak M.B., Durrell P.R. Astrophys. J. 577, 31 (2002) Planetary nebulae as standard candles. XII. Connecting the population I and population II distance scales.

J/ApJ/577/738/  table2.dat table4.dat table5.dat  (36Kb)

Kong A.K.H., Garcia M.R., Primini F.A., Murray S.S., di Stefano R., McClintock J.E. Astrophys. J. 577, 738 (2002) X-ray point sources in the central region of M31 as seen by Chandra.

J/ApJ/578/114/  table1.dat table2.dat  (16Kb)

Kaaret P. Astrophys. J. 578, 114 (2002) A Chandra High Resolution Camera observation of X-ray point sources in M31.

J/ApJ/578/405/  table1.dat  (9Kb)

Rutledge R.E., Bildsten L., Brown E.F., Pavlov G.G., Zavlin V.E. Astrophys. J. 578, 405 (2002) A possible transient neutron star in quiescence in the globular cluster NGC 5139.

J/ApJ/578/914/  table1.dat  (13Kb)

Hartmann L. Astrophys. J. 578, 914 (2002) Flows, fragmentation, and star formation. I. Low-mass stars in Taurus.

J/ApJ/578/943/  refs.dat table2.dat  (19Kb)

de Medeiros J.R., da Silva J.R.P., Maia M.R.G. Astrophys. J. 578, 943 (2002) The rotation of binary systems with evolved components.

J/ApJ/579/93/  table2.dat table4.dat table5.dat  (75Kb)

Postman M., Lauer T.R., Oegerle W., Donahue M. Astrophys. J. 579, 93 (2002) The KPNO/Deeprange distant cluster survey. I. The catalog and the space density of intermediate-redshift clusters.

J/ApJ/580/774/  refs.dat table1.dat  (18Kb)

Ebeling H., Mullis C.R., Tully R.B. Astrophys. J. 580, 774 (2002) A systematic X-ray search for clusters of galaxies behind the Milky way.

J/ApJ/581/258/  table3.dat  (23Kb)

Figer D.F., Najarro F., Gilmore D., Morris M., Kim S.S., Serabyn E., McLean I.S., Gilbert A.M., Graham J.R., Larkin J.E., Levenson N.A., Teplitz H.I.

Astrophys. J. 581, 258 (2002) Massive stars in the Arches Cluster.

J/ApJ/581/325/  table2.dat table3.dat  (32Kb)

Seth A.C., Greenhill L.J., Holder B.P. Astrophys. J. 581, 325 (2002) Water masers as tracers of protostellar disks and outflows in the intermediate-mass star-forming region NGC 2071.

J/ApJ/581/844/  table1.dat  (16Kb)

Smail I., Owen F.N., Morrison G.E., Keel W.C., Ivison R.J., Ledlow M.J. Astrophys. J., 581, 844-864 (2002) The diversity of Extremely Red Objects.

J/ApJ/581/1194/  catalog.dat  (22Kb)

Palla F., Stahler S.W. Astrophys. J. 581, 1194 (2002) Star formation in space and time: Taurus-Auriga.

J/ApJ/582/170/  table1.dat  (8Kb)

Durrell P.R., Mihos J.C., Feldmeier J.J., Jacoby G.H., Ciardullo R. Astrophys. J. 582, 170 (2003) Kinematics of planetary nebulae in M51’s tidal debris.

J/ApJ/582/382/  table1.dat table2.dat table4.dat  (114Kb)

Flaccomio E., Damiani F., Micela G., Sciortino S., Harnden F.R.Jr, Murray S.S., Wolk S.J. Astrophys. J. 582, 382 (2003) Chandra X-ray observation of the Orion Nebula Cluster. I. Detection, identification, and determination of X-ray luminosities.

J/ApJ/582/615/  table1.dat table2.dat table3.dat table5.dat  (27Kb)

Nandra K., Georgantopoulos I., Ptak A., Turner T.J. Astrophys. J. 582, 615 (2003) SHEEP: the search for the high-energy extragalactic population.

J/ApJ/582/668/  table1.dat  (20Kb)

Barton Gillespie E., Geller M.J., Kenyon S.J. Astrophys. J. 582, 668 (2003) Tidally triggered star formation in close pairs of galaxies. II. Constraints on burst strengths and ages.

J/ApJ/582/756/  table1.dat  (9Kb)

Kolpak M.A., Jackson J.M., Bania T.M., Clemens D.P., Dickey J.M. Astrophys. J. 582, 756 (2003) Resolving the kinematic ambiguity toward galactic H II regions.

J/ApJ/582/1011/  table2.dat  (10.2 Mb)

Salim S., Gould A. Astrophys. J. 582, 1011 (2003) Improved astrometry and photometry for the Luyten catalog. II. Faint stars and the revised catalog.

J/ApJ/582/1123/  stars.dat table1.dat  (24Kb)

Mallen-Ornelas G., Seager S., Yee H.K.C., Minniti D., Gladders M.D., Mallen-Fullerton G.M., Brown T.M. Astrophys. J. 582, 1123 (2003) The EXPLORE project. I. A deep search for transiting extrasolar planets.

J/ApJ/584/911/  table1.dat  (60Kb)

Feigelson E.D., Gaffney III J.A., Garmire G., Hillenbrand L.A., Townsley L. Astrophys. J. 584, 911 (2003) X-rays in the Orion nebula cluster: constraints on the origins of magnetic activity in pre-main-sequence stars.

J/ApJ/585/647/  table1.dat  (12Kb)

Boroson T.A. Astrophys. J. 585, 647 (2003) Does the narrow [O III] λ5007 line reflect the velocity dispersion in active galactic nuclei?

J/ApJ/586/765/  table1.dat table2.dat  (19Kb)

Yan L., Thompson D. Astrophys. J. 586, 765 (2003) Hubble Space Telescope WFPC2 morphologies of K-selected extremely red galaxies.

J/ApJ/586/794/  refs.dat tablea1.dat  (42Kb)

Bell E.F.

Astrophys. J. 586, 794 (2003) Estimating star formation rates from infrared and radio luminosities: the origin of the radio-infrared correlation.

J/ApJ/586/814/  table1.dat table2.dat table3.dat  (106Kb)

Maccarone T.J., Kundu A., Zepf S.E. Astrophys. J. 586, 814 (2003) The low-mass X-ray binary-globular connection. II. NGC 4472 X-ray source properties and source catalogs.

J/ApJ/586/826/  table1.dat  (14Kb)

Kim D.-W., Fabbiano G. Astrophys. J. 586, 826 (2003) Chandra X-ray observations of NGC 1316 (Fornax A).

J/ApJ/586/1356/  table1.dat table2.dat  (16Kb)

Wachter S., Hoard D.W., Hansen K.H., Wilcox R.E., Taylor H.M., Finkelstein S.L. Astrophys. J. 586, 1356 (2003) Cool companions to white dwarfs from the Two Micron All-Sky Survey Second Incremental Data Release.

J/ApJ/588/218/  table1.dat  (118Kb)

Abraham R.G., van den Bergh S., Nair P. Astrophys. J. 588, 218 (2003) A new approach to galaxy morphology. I. Analysis of the Sloan Digital Sky Survey Early Data Release.

J/ApJ/588/824/  table1.dat  (43Kb)

Yanny B., Newberg H.J., Grebel E.K., Kent S., Odenkirchen M., Rockosi C.M., Schlegel D., Subbarao M., Brinkmann J., Fukugita M., Ivezic Z., Lamb D.Q., Schneider D.P., York D.G. Astrophys. J. 588, 824 (2003) A low-latitude halo stream around the Milky way.

J/ApJ/588/1009/  table4.dat table5.dat table6.dat  (40Kb)

Damiani F., Flaccomio E., Micela G., Sciortino S., Harnden F.R.Jr, Murray S.S., Wolk S.J., Jeffries R.D. Astrophys. J. 588, 1009 (2003) Chandra X-ray observations of the young open cluster NGC 2516.

J/ApJ/589/225/  table3.dat  (653Kb)

Muno M.P., Baganoff F.K., Bautz M.W., Brandt W.N., Broos P.S., Feigelson E.D., Garmire G.P., Morris M.R., Ricker G.R., Townsley L.K. Astrophys. J. 589, 225 (2003) A deep Chandra catalog of X-ray point sources toward the Galactic Center.

J/ApJ/589/509/  tab12.dat tab34.dat tab68.dat table5.dat table7.dat  (47Kb)

Evans N.R., Seward F.D., Krauss M.I., Isobe T., Nichols J., Schlegel E.M., Wolk S.J. Astrophys. J. 589, 509 (2003) Chandra observations of associates of ηCarinae. I. Luminosities.

J/ApJ/589/733/  table1.dat  (23Kb)

Pollack L.K., Taylor G.B., Zavala R.T. Astrophys. J. 589, 733 (2003) VLBI polarimetry of 177 sources from the Caltech-Jodrell Bank Flat-Spectrum Survey.

J/ApJ/590/73/  table1.dat  (24Kb)

Xu D.W., Komossa S., Wei J.Y., Qian Y., Zheng X.Z. Astrophys. J. 590, 73 (2003) An active galactic nucleus sample with high X-Ray-to-Optical flux ratio from RASS. II. Optical emission line properties of Seyfert 1-Type active galactic nuclei.

J/ApJ/590/109/  notes.dat table1.dat table2.dat  (35Kb)

Sowards-Emmerd D., Romani R.W., Michelson P.F. Astrophys. J., 590, 109-122 (2003) The gamma-ray blazar content of the northern sky.

J/ApJ/590/797/  table1.dat  (9Kb)

Van Den Bergh S. Astrophys. J., 590, 797-802 (2003) Some global characteristics of the galactic globular cluster system.

J/ApJ/591/L111/  table1.dat  (12Kb)

Bonanos A.Z., Stanek K.Z. Astrophys. J. 591, L111 (2003) Reanalysis of Very Large Telescope data for M83 with image subtraction-Ninefold increase in number of

Cepheids.

J/ApJ/591/53/  table4.dat  (2.4 Mb)

Treu T., Ellis R.S., Kneib J.-P., Dressler A., Smail I., Czoske O., Oemler A., Natarajan P. Astrophys. J. 591, 53 (2003) A Wide-Field Hubble Space Telescope study of the cluster Cl 0024+16 at z = 0.4. I. Morphological distributions to 5 Mpc radius.

J/ApJ/591/138/  table1.dat  (15Kb)

Smith D.A., Wilson A.S. Astrophys. J., 591, 138-153 (2003) A Chandra X-ray study of NGC 1068. II. The luminous X-ray source population.

J/ApJ/591/640/  refs.dat table2.dat table3.dat table4.dat  (73Kb)

Cohen A.S., Rottgering H.J.A., Kassim N.E., Cotton W.D., Perley R.A., Wilman R., Best P., Pierre M., Birkinshaw M., Bremer M., Zanichelli A. Astrophys. J. 591, 640 (2003) The low-frequency radio counterpart of the XMM Large-Scale Structure Survey.

J/ApJ/591/764/  table2.dat  (238Kb)

Christlein D., Zabludoff A.I. Astrophys. J. 591, 764 (2003) Galaxy luminosity functions from deep spectroscopic samples of rich clusters.

J/ApJ/591/850/  table2.dat  (36Kb)

Cote P., McLaughlin D.E., Cohen J.G., Blakeslee J.P. Astrophys. J. 591, 850 (2003) Dynamics of the globular cluster system associated with M49 (NGC 4472): cluster orbital properties and the distribution of dark matter.

J/ApJ/592/728/ table1.dat table7.dat table8.dat table9.dat table10.dat table11.dat table12.dat table13.dat table14.dat table15.dat table16.dat table17.dat table18.dat table19.dat table20.dat table21.dat table22.dat table23.dat (232Kb) Steidel C.C., Adelberger K.L., Shapley A.E., Pettini M., Dickinson M., Giavalisco M. Astrophys. J. 592, 728 (2003) Lyman break galaxies at redshift z 3: survey description and full data set.

J/ApJ/593/1093/  table2.dat table3.dat table4.dat table6.dat  (83Kb)

Luhman K.L., Stauffer J.R., Muench A.A., Rieke G.H., Lada E.A., Bouvier J., Lada C.J. Astrophys. J. 593, 1093 (2003) A census of the young cluster IC 348.

J/ApJ/594/1/  refs.dat table4.dat table5.dat table6.dat table8.dat table12.dat  (53Kb)

Tonry J.L., Schmidt B.P., Barris B., Candia P., Challis P., Clocchiatti A., Coil A.L., Filippenko A.V., Garnavich P., Hogan C., Holland S.T., Jha S., Kirshner R.P., Krisciunas K., Leibundgut B., Li W., Matheson T., Phillips M.M., Riess A.G., Schommer R., Smith R.C., Sollerman J., Spyromilio J., Stubbs C.W., Suntzeff N.B. Astrophys. J. 594, 1 (2003) Cosmological results from high-z supernovae.

J/ApJ/594/154/  refs.dat table4.dat  (36Kb)

Mullis C.R., McNamara B.R., Quintana H., Vikhlinin A., Henry J.P., Gioia I.M., Hornstrup A., Forman W., Jones C. Astrophys. J. 594, 154 (2003) The 160 square degree ROSAT survey: the revised catalog of 201 clusters with spectroscopic redshifts.

J/ApJ/595/187/  table2.dat  (106Kb)

McNamara B.J., Harrison T.E., Anderson J. Astrophys. J. 595, 187 (2003) Does M15 possess an intermediate-mass black hole in its core?

J/ApJ/595/247/  table2.dat  (17Kb)

Ruiz M.T., Peimbert A., Peimbert M., Esteban C. Astrophys. J. 595, 247 (2003) Very Large Telescope echelle spectrophotometry of the planetary nebula NGC 5307 and temperature variations.

J/ApJ/595/685/  table1.dat table2.dat  (221Kb)

Grogin N.A., Koekemoer A.M., Schreier E.J., Bergeron J., Giacconi R., Hasinger G., Kewley L., Norman C., Rosati P., Tozzi P., Zirm A. Astrophys. J. 595, 685 (2003) Hubble Space Telescope imaging in the Chandra Deep Field-South. III. Quantitative morphology of the 1

million second Chandra counterparts and comparison with the field population.

J/ApJ/595/1154/ table3.dat table4.dat table5.dat table6.dat table7.dat table9.dat table10.dat table11.dat table12.dat (60Kb) Fulbright J.P., Johnson J.A. Astrophys. J. 595, 1154 (2003) Oxygen abundances in metal-poor stars.

J/ApJ/595/1206/  table1.dat table4.dat table5.dat  (213Kb)

Suchkov A.A., Makarov V.V., Voges W. Astrophys. J. 595, 1206 (2003) ROSAT view of Hipparcos F stars.

J/ApJ/596/437/  table2.dat table3.dat  (20Kb)

Chisholm J.R., Dodelson S., Kolb E.W. Astrophys. J., 596, 437-450 (2003) Stellar-mass black holes in the solar neighborhood.

J/ApJ/596/748/  table2.dat table3.dat  (228Kb)

Thompson R.I. Astrophys. J., 596, 748-767 (2003) Star formation history and other properties of the Northern Hubble Deep Field.

J/ApJ/596/768/  table1.dat table4.dat table5.dat  (63Kb)

Schaye J., Aguirre A., Kim T.-S., Theuns T., Rauch M., Sargent W.L.W. Astrophys. J., 596, 768-796 (2003) Metallicity of the intergalactic medium using pixel statistics. II. The distribution of metals as traced by C IV.

J/ApJ/596/944/  table2.dat table4.dat table6.dat  (220Kb)

Harrison F.A., Eckart M.E., Mao P.H., Helfand D.J., Stern D. Astrophys. J., 596, 944-956 (2003) The Serendipitous Extragalactic X-ray Source Identification program. I. Characteristics of the hard X-ray sample.

J/ApJ/596/1064/  table1.dat  (29Kb)

Jiang Z., Yao Y., Yang J., Baba D., Kato D., Kurita M., Nagashima C., Nagata T., Nagayama T., Nakajima Y., Ishii M., Tamura M., Sugitani K. Astrophys. J., 596, 1064-1079 (2003) A near-infrared study of the star-forming region S269.

J/ApJ/597/323/  refs.dat table2.dat table4.dat table6.dat table13.dat  (46Kb)

Blum R.D., Ramirez S.V., Sellgren K., Olsen K. Astrophys. J., 597, 323-346 (2003) Really cool stars and the star formation history at the Galactic Center.

J/ApJ/597/555/  table4.dat  (22Kb)

Mainzer A.K., Mclean I.S. Astrophys. J., 597, 555-565 (2003) Using narrowband photometry to detect young brown dwarfs in IC 348.

J/ApJ/598/L43/  table1.dat table2.dat  (11Kb)

Wisniewski J.P., Bjorkman K.S., Magalhaes A.M. Astrophys. J., 598, L43-L46 (2003) Evolution of the inner circumstellar envelope of V838 Monocerotis.

J/ApJ/598/260/  table2.dat table3.dat  (248Kb)

Prada F., Vitvitska M., Klypin A., Holtzman J.A., Schlegel D.J., Grebel E.K., Rix H.-W., Brinkmann J., Mckay T.A., Csabai I. Astrophys. J., 598, 260-271 (2003) Observing the dark matter density profile of isolated galaxies.

J/ApJ/598/375/  notes.dat table1.dat  (60Kb)

Skinner S., Gagne M., Belzer E. Astrophys. J., 598, 375-391 (2003) A deep Chandra X-ray observation of the embedded young cluster in NGC 2024.

J/ApJ/598/597/  table2.dat table3.dat table4.dat table5.dat  (624Kb)

Alcock C., Alves D.R., Becker A., Bennett D., Cook K.H., Drake A., Freeman K., Geha M., Griest K., Kovacs G., Lehner M., Marshall S., Minniti D., Nelson C., Peterson B., Popowski P., Pratt M., Quinn P., Rodgers A., Stubbs C., Sutherland W., Vandehei T., Welch D.L. (the Macho Collaboration) Astrophys. J., 598, 597-609 (2003)

The MACHO project Large Magellanic Cloud variable star inventory. XI. Frequency analysis of the fundamental-mode RR Lyrae stars.

J/ApJ/598/969/  table2.dat  (12Kb)

Wang Q.D., Chaves T., Irwin J.A. Astrophys. J., 598, 969-981 (2003) Chandra observation of the edge-on galaxy NGC 3556 (M108): violent galactic disk-halo interaction revealed.

J/ApJ/599/116/  table1.dat table2.dat  (16Kb)

Vestergaard M. Astrophys. J., 599, 116-139 (2003) Occurrence and global properties of narrow C IV λ1549Åabsorption lines in moderate-redshift quasars.

J/ApJ/599/218/  table1.dat table2.dat  (24Kb)

Sivakoff G.R., Sarazin C.L., Irwin J.A. Astrophys. J., 599, 218-236 (2003) Chandra observations of low-mass X-ray binaries and diffuse gas in the early-type galaxies NGC 4365 and NGC 4382 (M85).

J/ApJ/599/886/  refs.dat table5.dat  (14Kb)

Eracleous M., Halpern J.P. Astrophys. J., 599, 886-908 (2003) Completion of a survey and detailed study of double-peaked emission lines in radio-loud active galactic nuclei.

J/ApJ/599/1006/  table1.dat  (19Kb)

Kobulnicky H.A., Willmer C.N.A., Phillips A.C., Koo D.C., Faber S.M., Weiner B.J., Sarajedini V.L., Simard L., Vogt N.P. Astrophys. J., 599, 1006-1030 (2003); Erratum Vol. 610, p. 1234 (2004) The DEEP Groth Strip Survey. VII. The metallicity of field galaxies at 0.26 z 0.82 and the evolution of the luminosity-metallicity relation.

J/ApJ/599/1049/  table2.dat  (9Kb)

Wilson C.D., Scoville N., Madden S.C., Charmandaris V. Astrophys. J., 599, 1049-1066 (2003) The mass function of supergiant molecular complexes and implications for forming young massive star clusters in the Antennae (NGC 4038/4039).

J/ApJ/599/1067/  table1.dat  (16Kb)

Di Stefano R., Kong A.K.H., Vandalfsen M.L., Harris W.E., Murray S.S., Delain K.M. Astrophys. J., 599, 1067-1081 (2003) X-ray point sources in the Sombrero galaxy: very soft sources, the globular cluster/low-mass X-ray binary connection, and an overview.

J/ApJ/600/L155/  table1.dat table2.dat  (11Kb)

Fassnacht C.D., Moustakas L.A., Casertano S., Ferguson H.C., Lucas R.A., Park Y. Astrophys. J., 600, L155-L158 (2004) Strong gravitational lens candidates in the GOODS ACS fields.

J/ApJ/600/L93 See II/261/

J/ApJ/600/L99/  table1.dat  (33Kb)

Dickinson M., Stern D., Giavalisco M., Ferguson H.C., Tsvetanov Z., Chornock R., Cristiani S., Dawson S., Dey A., Filippenko A.V., Moustakas L.A., Nonino M., Papovich C., Ravindranath S., Riess A., Rosati P., Spinrad H., Vanzella E. Astrophys. J., 600, L99-L102 (2004) Color-selected galaxies at z 6 in the Great Observatories Origins Deep Survey.

J/ApJ/600/716/  table1.dat table2.dat table3.dat  (22Kb)

Minniti D., Rejkuba M., Funes J.G., Akiyama S. Astrophys. J., 600, 716-728 (2004) Optical counterparts of X-ray point sources observed by Chandra in NGC 5128: 20 new globular cluster X-ray sources.

J/ApJ/600/729/  table1.dat  (24Kb)

Randall S.W., Sarazin C.L., Irwin J.A. Astrophys. J., 600, 729-742 (2004) Chandra observation of diffuse gas and low-mass X-ray binaries in the elliptical galaxy NGC 4649 (M60).

J/ApJ/601/485/  table1a.dat table1b.dat table1c.dat table1d.dat table2.dat  (27Kb)

Franchini M., Morossi C., Di Marcantonio P., Malagnini M.L., Chavez M., Rodriguez-Merino L. Astrophys. J., 601, 485-499 (2004)

Synthetic Lick indices and detection of α-enhanced stars.

J/ApJ/601/735/  table2.dat table3.dat table4.dat  (25Kb)

Terashima Y., Wilson A.S. Astrophys. J., 601, 735-758 (2004) The luminous X-ray source population in M51 observed with Chandra.

J/ApJ/601/979/  refs.dat table1.dat table2.dat  (44Kb)

Wolff S.C., Strom S.E., Hillenbrand L.A. Astrophys. J., 601, 979-999 (2004) The angular momentum evolution of 0.1-10M_{⊙} stars from the birth line to the main sequence.

J/ApJ/602/L117/  table1.dat  (8Kb)

Schuler S.C., King J.R., Hobbs L.M., Pinsonneault M.H. Astrophys. J., 602, L117-L120 (2004) Oxygen in open cluster dwarfs: Pleiades and M34.

J/ApJ/602/200/  refs.dat table1.dat table3.dat table4.dat table5.dat  (211Kb)

Izotov Y.I., Thuan T.X. Astrophys. J., 602, 200-230 (2004) Systematic effects and a new determination of the primordial abundance of ^{4}He and dY/dZ from observations of blue compact galaxies.

J/ApJ/602/231/  table1.dat table2.dat tablea1.dat  (152Kb)

Colbert E.J.M., Heckman T.M., Ptak A.F., Strickland D.K., Weaver K.A. Astrophys. J., 602, 231-248 (2004) Old and young X-ray point source populations in nearby galaxies.

J/ApJ/602/264/  table1.dat  (19Kb)

Mcnamara B.J., Harrison T.E., Baumgardt H. Astrophys. J., 602, 264-270 (2004) The dynamical distance to M15: estimates of the cluster’s age and mass and of the absolute magnitude of its RR Lyrae stars.

J/ApJ/602/571/  table1.dat table4.dat table5.dat table6.dat table8.dat  (59Kb)

Barris B.J., Tonry J.L., Blondin S., Challis P., Chornock R., Clocchiatti A., Filippenko A.V., Garnavich P., Holland S.T., Jha S., Kirshner R.P., Krisciunas K., Leibundgut B., Li W., Matheson T., Miknaitis G., Riess A.G., Schmidt B.P., Smith R.C., Sollerman J., Spyromilio J., Stubbs C.W., Suntzeff N.B., Aussel H., Chambers K.C., Connelley M.S., Donovan D., Henry J.P., Kaiser N., Liu M.C., Martin E.L., Wainscoat R.J. Astrophys. J., 602, 571-594 (2004) Twenty-three high-redshift supernovae from the Institute for Astronomy Deep Survey: doubling the supernova sample at z 0.7.

J/ApJ/602/664/  table1.dat  (24Kb)

Gutierrez C.M., Trujillo I., Aguerri J.A.L., Graham A.W., Caon N. Astrophys. J., 602, 664-677 (2004) Quantitative morphology of galaxies in the core of the Coma cluster.

J/ApJ/602/685/  table2.dat table3.dat table6.dat table7.dat  (74Kb)

Peng E.W., Ford H.C., Freeman K.C. Astrophys. J., 602, 685-704 (2004) The planetary nebula system and dynamics in the outer halo of NGC 5128.

J/ApJ/602/850/  table1.dat  (14Kb)

Shepherd D.S., Nurnberger D.E.A., Bronfman L. Astrophys. J., 602, 850-859 (2004) Discovery of a massive protostar near IRAS 18507+0121.

J/ApJ/603/503/  table2.dat table3.dat table4.dat table5.dat table6.dat  (65Kb)

Harris J., Calzetti D., Gallagher Iii J.S., Smith D.A., Conselice C.J. Astrophys. J., 603, 503-522 (2004) The recent cluster formation histories of NGC 5253 and NGC 3077: environmental impact on star formation.

J/ApJ/603/584/  notes.dat table2.dat table3.dat table5.dat  (169Kb)

Pereyra A., Magalhaes A.M. Astrophys. J., 603, 584-594 (2004) Polarimetry toward the Musca dark cloud. I. The catalog.

J/ApJ/604/141/  table2.dat  (16Kb)

Finn R.A., Zaritsky D., Mccarthy D.W.Jr Astrophys. J., 604, 141-152 (2004)

Hα-derived star formation rates for the z=0.84 galaxy cluster Cl J0023+0423B.

J/ApJ/604/521/  table1.dat  (16Kb)

Trujillo I., Rudnick G., Rix H.-W., Labbe I., Franx M., Daddi E., Van Dokkum P.G., Forster Schreiber N.M., Kuijken K., Moorwood A., Rottgering H., Van De Wel A., Van Der Werf P., Van Starkenburg L. Astrophys. J., 604, 521-533 (2004) The luminosity-size and mass-size relations of galaxies out to z 3.

J/ApJ/604/717/  table2.dat table3.dat table4.dat table5.dat  (35Kb)

Houde M., Dowell C.D., Hildebrand R.H., Dotson J.L., Vaillancourt J.E., Phillips T.G., Peng R., Bastien P. Astrophys. J., 604, 717-740 (2004) Tracing the magnetic field in Orion A.

J/ApJ/605/259/  table1.dat  (14Kb)

Pravdo S.H., Tsuboi Y., Maeda Y. Astrophys. J., 605, 259-271 (2004) X-rays from HH 80, HH 81, and the central region.

J/ApJ/605/759/  position.dat table2.dat  (31Kb)

Bhat N.D.R., Cordes J.M., Camilo F., Nice D.J., Lorimer D.R. Astrophys. J., 605, 759-783 (2004) Multifrequency observations of radio pulse broadening and constraints on interstellar electron density microstructure.

J/ApJ/606/L25/  table1.dat  (9Kb)

Bouwens R.J., Illingworth G.D., Thompson R.I., Blakeslee J.P., Dickinson M.E., Broadhurst T.J., Eisenstein D.J., Fan X., Franx M., Meurer G., Van Dokkum P. Astrophys. J., 606, L25-L28 (2004) Star formation at z 6: the Hubble Ultra Deep Parallel Fields.

J/ApJ/606/L73/  table1.dat  (16Kb)

Blake G.A., Boogert A.C.A. Astrophys. J., 606, L73-L76 (2004) High-resolution 4.7 micron Keck/NIRSPEC spectroscopy of the CO emission from the disks surrounding Herbig Ae stars.

J/ApJ/606/466/  table1.dat  (14Kb)

Wolk S.J., Harnden F.R.JR, Murray S.S., Adams N.R., Damiani F., Flaccomio E., Micela G., Sciortino S., Jeffries R.D. Astrophys. J., 606, 466-482 (2004) Coronal variability in the young cluster NGC 2516.

J/ApJ/606/664/  table1.dat table2.dat  (23Kb)

Dannerbauer H., Lehnert M.D., Lutz D., Tacconi L., Bertoldi F., Carilli C., Genzel R., Menten K.M. Astrophys. J., 606, 664-682 (2004) The faint counterparts of MAMBO millimeter sources near the New Technology Telescope Deep Field.

J/ApJ/607/60/  table2.dat  (14Kb)

Glikman E., Gregg M.D., Lacy M., Helfand D.J., Becker R.H., White R.L. Astrophys. J., 607, 60-75 (2004) FIRST-2MASS sources below the APM detection threshold: a population of highly reddened quasars.

J/ApJ/607/202/  table1.dat table2.dat  (138Kb)

Mahdavi A., Geller M.J. Astrophys. J. 607, 202 (2004) A redshift survey of nearby galaxy groups: the shape of the mass density profile

J/ApJ/607/426/  table3.dat table4.dat table5.dat table6.dat table11.dat table12.dat  (935Kb)

Kleinman S.J., Harris H.C., Eisenstein D.J., Liebert J., Nitta A., Krzesinski J., Munn J.A., Dahn C.C., Hawley S.L., Pier J.R., Schmidt G., Silvestri N.M., Smith J.A., Szkody P., Strauss M.A., Knapp G.R., Collinge M.J., Mukadam A.S., Koester D., Uomoto A., Schlegel D.J., Anderson S.F., Brinkmann J., Lamb D.Q., Schneider D.P., York D.G. Astrophys. J., 607, 426-444 (2004) A catalog of spectroscopically identified white dwarf stars in the First Data Release of the Sloan Digital Sky Survey.

J/ApJ/607/665/  table1.dat table2.dat table5.dat  (28Kb)

Riess A.G., Strolger L.-G., Tonry J., Casertano S., Ferguson H.C., Mobasher B., Challis P., Filippenko A.V., Jha S., Li W., Chornock R., Kirshner R.P., Leibundgut B., Dickinson M., Livio M., Giavalisco M., Steidel C.C., Benitez T., Tsvetanov Z. Astrophys. J., 607, 665-687 (2004)

Type Ia supernova discoveries at z 1 from the Hubble Space Telescope: evidence for past deceleration and constraints on dark energy evolution.

J/ApJ/607/904/  table1.dat  (21Kb)

Rho J., Ramirez S.V., Corcoran M.F., Hamaguchi K., Lefloch B. Astrophys. J., 607, 904-912 (2004) Chandra observation of the Trifid nebula: X-ray emission from the O star complex and actively forming pre-main-sequence stars.

J/ApJ/608/L93/  table1.dat  (6Kb)

Cobb B.E., Bailyn C.D., Van Dokkum P.G., Buxton M.M., Bloom J.S. Astrophys. J., 608, L93-L96 (2004) The supernova associated with GRB 031203: SMARTS optical-infrared light curves from 0.2 to 92 days.

J/ApJ/608/243/  photom.dat  (1.5 Mb)

Hargis J.R., Sandquist E.L., Bolte M. Astrophys. J., 608, 243-260 (2004) The luminosity function and color-magnitude diagram of the globular cluster M12.

J/ApJ/608/405/  yields.dat  (75Kb)

Chieffi A., Limongi M. Astrophys. J., 608, 405-410 (2004) Explosive yields of massive stars from z=0 to z=z_{⊙}.

J/ApJ/608/781/  table1.dat  (131Kb)

Damiani F., Flaccomio E., Micela G., Sciortino S., Harnden F.R.Jr, Murray S.S. Astrophys. J., 608, 781-796 (2004) A deep Chandra X-ray observation of the rich young cluster NGC 6530. I. The X-ray source catalog and the cluster population.

J/ApJ/609/539/  table1.dat table2.dat  (38Kb)

Kellermann K.I., Lister M.L., Homan D.C., Vermeulen R.C., Cohen M.H., Ros E., Kadler M., Zensus J.A., Kovalev Y.Y. Astrophys. J., 609, 539-563 (2004) Sub-milliarcsecond imaging of quasars and active galactic nuclei. III. Kinematics of parsec-scale radio jets.

J/ApJ/609/564/  table1.dat table2.dat  (31Kb)

Sowards-Emmerd D., Romani R.W., Michelson P.F., Ulvestad J.S. Astrophys. J., 609, 564-575 (2004) Blazar counterparts for 3EG sources at -40^{∘}Decl. 0^{∘}: pushing south through the bulge.

J/ApJ/609/735/  notes.dat table1.dat table2.dat  (33Kb)

Williams B.F., Garcia M.R., Kong A.K.H., Primini F.A., King A.R., Di Stefano R., Murray S.S. Astrophys. J., 609, 735-754 (2004) A synoptic X-ray study of M31 with the Chandra high resolution camera.

J/ApJ/609/935/  table2.dat  (51Kb)

Yonetoku D., Murakami T., Nakamura T., Yamazaki R., Inoue A.K., Ioka K. Astrophys. J., 609, 935-951 (2004) Gamma-ray burst formation rate inferred from the spectral peak energy-peak luminosity relation.

J/ApJ/610/128/  table1.dat  (22Kb)

Watanabe C., Ohta K., Akiyama M., Ueda Y. Astrophys. J., 610, 128-139 (2004) Near-infrared colors of hard X-ray-selected active galactic nuclei.

J/ApJ/610/247/  table1.dat  (12Kb)

Di Stefano R., Kong A.K.H., Greiner J., Primini F.A., Garcia M.R., Barmby P., Massey P., Hodge P.W., Williams B.F., Murray S.S., Curry S., Russo T.A. Astrophys. J., 610, 247-260 (2004) Supersoft X-ray sources in M31. I. A Chandra survey and an extension to quasi-soft sources.

J/ApJ/610/1064/  table1.dat table2.dat table3.dat  (37Kb)

Barrado Y Navascues D., Stauffer J.R., Bouvier J., Jayawardhana R., Cuillandre J.-C. Astrophys. J., 610, 1064-1078 (2004) The substellar population of the young cluster λOrionis.

J/ApJ/611/270/  table1.dat  (19Kb)

Gonzalez R.A., Liu M.C., Bruzual A.G. Astrophys. J., 611, 270-293 (2004) Infrared surface brightness fluctuations of Magellanic star clusters.

J/ApJ/611/323/  table1.dat table2.dat table3.dat table4.dat table5.dat  (296Kb)

Sandquist E.L., Bolte M. Astrophys. J., 611, 323-337 (2004) Exploring the upper red giant and asymptotic giant branches: the globular cluster M5.

J/ApJ/611/452/  table3.dat  (11Kb)

El Eid M.F., Meyer B.S., The L.-S. Astrophys. J., 611, 452-465 (2004) Evolution of massive stars up to the end of central oxygen burning.

J/ApJ/611/821/  table1.dat  (15Kb)

Wang Q.D., Owen F., Ledlow M. Astrophys. J., 611, 821-834 (2004) X-raying A2125: a large-scale hierarchical complex of galaxies and hot gas.

J/ApJ/612/L93/  table1.dat  (15Kb)

Yan H., Windhorst R.A. Astrophys. J., 612, L93-L96 (2004) Candidates of z 5.5-7 galaxies in the Hubble Space Telescope Ultra Deep Field.

J/ApJ/612/202/  table3.dat  (15Kb)

Menanteau F., Ford H.C., Illingworth G.D., Sirianni M., Blakeslee J.P., Meurer G.R., Martel A.R., Benitez N., Postman M., Franx M., Ardila D.R., Bartko F., Bouwens R.J., Broadhurst T.J., Brown R.A., Burrows C.J., Cheng E.S., Clampin M., Cross N.J.G., Feldman P.D., Golimowski D.A., Gronwall C., Hartig G.F., Infante L., Kimble R.A., Krist J.E., Lesser M.P., Miley G.K., Rosati P., Sparks W.B., Tran H.D., Tsvetanov Z.I., White R.L., Zheng W. Astrophys. J., 612, 202-214 (2004) Internal color properties of resolved spheroids in the deep Hubble Space Telescope advanced camera for surveys field of UGC 10214.

J/ApJ/612/437/  table2.dat table3.dat  (13Kb)

Suchkov A.A., Hanisch R.J. Astrophys. J., 612, 437-446 (2004) ClassX detection of candidate low-luminosity hard X-ray binaries.

J/ApJ/612/447/  table1.dat  (14Kb)

Fulbright J.P., Rich R.M., Castro S. Astrophys. J., 612, 447-453 (2004) Draco 119: a remarkable heavy-element-deficient giant.

J/ApJ/612/481/  table9.dat  (8Kb)

Proffitt C.R., Sahu K., Livio M., Krist J., Calzetti D., Gilliland R., Grady C., Lindler D., Woodgate B., Heap S., Clampin M., Gull T.R., Lisse C.M. Astrophys. J., 612, 481-495 (2004) Limits on the optical brightness of the ϵEridani dust ring.

J/ApJ/612/511/  table1.dat  (9Kb)

Lazio T.J.W., Farrell W.M., Dietrick J., Greenlees E., Hogan E., Jones C., Hennig L.A. Astrophys. J., 612, 511-518 (2004) The radiometric Bode’s law and extrasolar planets.

J/ApJ/612/652/  table1.dat table2.dat  (13Kb)

Daly R.A., Djorgovski S.G. Astrophys. J., 612, 652-659 (2004) Direct determination of the kinematics of the universe and properties of the dark energy as functions of redshift.

J/ApJ/612/848/  table1.dat  (12Kb)

Humphrey P.J., Buote D.A. Astrophys. J., 612, 848-859 (2004) A Chandra view of the normal S0 galaxy NGC 1332. I. An unbroken, steep power-law luminosity function for the low-mass X-ray binary population.

J/ApJ/613/129/  table2.dat  (53Kb)

Wise J.H., Eracleous M., Charlton J.C., Ganguly R. Astrophys. J., 613, 129-150 (2004) Variability of narrow, associated absorption lines in moderate- and low-redshift quasars.

J/ApJ/613/200/  table1.dat table2.dat  (16Kb)

Strolger L.-G., Riess A.G., Dahlen T., Livio M., Panagia N., Challis P., Tonry J.L., Filippenko A.V., Chornock R., Ferguson H., Koekemoer A., Mobasher B., Dickinson M., Giavalisco M., Casertano S., Hook R., Bondin S., Leibundgut B., Nonino M., Rosati P., Spinrad H., Steidel C.C., Stern D., Garnavich P.M., Matheson T., Grogin N., Hornschemeier A., Kretchmer C., Laidler V.G., Lee K., Lucas R., De Mello D., Moustakas L.A., Ravindranath S., Richardson M., Taylor E. Astrophys. J., 613, 200-223 (2004) The Hubble Higher z Supernova Search: supernovae to z 1.6 and constraints on type Ia progenitor models.

J/ApJ/613/279/  table1.dat  (17Kb)

Jordan A., Cote P., Ferrarese L., Blakeslee J.P., Mei S., Merritt D., Milosavljevic M., Peng E.W., Tonry J.L., West M.J. Astrophys. J., 613, 279-301 (2004) The ACS Virgo Cluster Survey. III. Chandra and Hubble space telescope observations of low-mass X-ray binaries and globular clusters in M87.

J/ApJ/613/312/  table2.dat  (27Kb)

Franchini M., Morossi C., Di Marcantonio P., Malagnini M.L., Chavez M., Rodriguez-Merino L. Astrophys. J., 613, 312-321 (2004) Synthetic Lick indices and detection of α-enhanced stars. II. F, G, and K stars in the -1.0 [Fe/H] +0.50 range.

J/ApJ/613/343/  table1.dat  (20Kb)

Delaney T., Rudnick L., Fesen R.A., Jones T.W., Petre R., Morse J.A. Astrophys. J., 613, 343-348 (2004) Kinematics of X-Ray-emitting components in Cassiopeia A.

J/ApJ/613/393/  notes.dat table1.dat table3.dat table4.dat  (35Kb)

Gagne M., Skinner S.L., Daniel K.J. Astrophys. J., 613, 393-415 (2004) Simultaneous Chandra and very large array observations of young stars and protostars in ρOphiuchus Cloud Core A.

J/ApJ/613/682/  refs.dat table1.dat table2.dat table5.dat table6.dat table8.dat  (60Kb)

Peterson B.M., Ferrarese L., Gilbert K.M., Kaspi S., Malkan M.A., Maoz D., Merritt D., Netzer H., Onken C.A., Pogge R.W., Vestergaard M., Wandel A. Astrophys. J., 613, 682-699 (2004) Central masses and broad-line region sizes of active galactic nuclei. II. A homogeneous analysis of a large reverberation-mapping database.

J/ApJ/613/841/  table2.dat table3.dat  (54Kb)

Miller N.A., Owen F.N., Hill J.M., Keel W.C., Ledlow M.J., Oegerle W.R. Astrophys. J., 613, 841-850 (2004) The dynamics of A2125.

J/ApJ/613/914/  table1.dat table4.dat table5.dat table6.dat  (33Kb)

Helmboldt J.F., Walterbos R.A.M., Bothun G.D., O’Neil K., De Blok W.J.G. Astrophys. J., 613, 914-932 (2004) Star formation in HI-selected galaxies. I. Sample characteristics.

J/ApJ/613/1179/  table2.dat table3.dat table6.dat table7.dat  (115Kb)

Muno M.P., Arabadjis J.S., Baganoff F.K., Bautz M.W., Brandt W.N., Broos P.S., Feigelson E.D., Garmire G.P., Morris M.R., Ricker G.R. Astrophys. J., 613, 1179-1201 (2004) The spectra and variability of X-ray sources in a deep Chandra observation of the Galactic Center.

J/ApJ/614/75/  table3.dat table4.dat  (11Kb)

Francis P.J., Palunas P., Teplitz H.I., Williger G.M., Woodgate B.E. Astrophys. J., 614, 75-83 (2004) The distribution of Lyα-emitting galaxies at z=2.38. II. Spectroscopy.

J/ApJ/614/91/  table1.dat table2.dat  (17Kb)

Wu X.-B., Liu F.K. Astrophys. J., 614, 91-100 (2004) Black hole mass and accretion rate of active galactic nuclei with double-peaked broad emission lines.

J/ApJ/614/167/  table2.dat  (17Kb)

Ciardullo R., Durrell P.R., Laychak M.B., Herrmann K.A., Moody K., Jacoby G.H., Feldmeier J.J. Astrophys. J., 614, 167-185 (2004) The planetary nebula system of M33.

J/ApJ/614/267/  table1.dat table2.dat table3.dat  (22Kb)

Feigelson E.D., Lawson W.A. Astrophys. J., 614, 267-283 (2004) An X-ray census of young stars in the Chamaeleon I North cloud.

J/ApJ/614/607/  table1.dat table2.dat  (10Kb)

Ojha R., Fey A.L., Jauncey D.L., Lovell J.E.J., Johnston K.J. Astrophys. J., 614, 607-614 (2004) Milliarcsecond structure of microarcsecond sources: comparison of scintillating and nonscintillating extragalactic radio sources.

J/ApJ/614/671/  table1.dat  (5Kb)

Chapman S.C., Smail I., Blain A.W., Ivison R.J. Astrophys. J., 614, 671-678 (2004) A population of hot, dusty ultraluminous galaxies at z 2.

J/ApJ/614/745/  table2.dat table4.dat  (18Kb)

Hyung S., Feibelman W.A. Astrophys. J., 614, 745-756 (2004) Optical and IUE spectra of the planetary nebula NGC 7026.

J/ApJ/615/702/  table3.dat table4.dat table5.dat  (33Kb)

Argon A.L., Greenhill L.J., Moran J.M., Reid M.J., Menten K.M., Inoue M. Astrophys. J., 615, 702-719 (2004) The IC 133 water vapor maser in the galaxy M33: a geometric distance.

J/ApJ/615/887/  table2.dat table3.dat table4.dat table5.dat  (28Kb)

Wachter S., Patel S.K., Kouveliotou C., Bouchet P., Ozel F., Tennant A.F., Woods P.M., Hurley K., Becker W., Slane P. Astrophys. J., 615, 887-896 (2004) Precise localization of the soft gamma repeater SGR 1627-41 and the anomalous X-ray pulsar AXP 1E1841-045 with Chandra.

J/ApJ/615/897/  table1.dat table2.dat  (13Kb)

Becker W., Weisskopf M.C., Arzoumanian Z., Lorimer D., Camilo F., Elsner R.F., Kanbach G., Reimer O., Swartz D.A., Tennant A.F., O’Dell S.L. Astrophys. J., 615, 897-907 (2004) A multiwavelength search for a counterpart of the brightest unidentified gamma-ray source 3EG J2020+4017 (2CG 078+2).

J/ApJ/616/71/  table1.dat  (11Kb)

Smail I., Chapman S.C., Blain A.W., Ivison R.J. Astrophys. J., 616, 71-85 (2004) The rest-frame optical properties of SCUBA galaxies.

J/ApJ/616/110/  table1.dat  (26Kb)

Horiuchi S., Fomalont E.B., Scott W.K., Taylor A.R., Lovell J.E.J., Moellenbrock G.A., Dodson R., Murata Y., Hirabayashi H., Edwards P.G., Gurvits L.I., Shen Z.-Q. Astrophys. J., 616, 110-122 (2004) The VSOP 5 GHz active galactic nucleus survey. IV. The angular Size/Brightness temperature distribution.

J/ApJ/616/562/  table1.dat table2.dat table3.dat  (19Kb)

Abt H.A., Boonyarak C. Astrophys. J., 616, 562-566 (2004) Tidal effects in binaries of various periods.

J/ApJ/616/707/  table2.dat table4.dat  (30Kb)

Hunt L.K., Malkan M.A. Astrophys. J., 616, 707-729 (2004) Circumnuclear structure and black hole fueling: Hubble Space Telescope NICMOS imaging of 250 active and normal galaxies.

J/ApJ/616/804/  table1.dat  (10Kb)

Hurley-Keller D., Morrison H.L., Harding P., Jacoby G.H. Astrophys. J., 616, 804-820 (2004) Planetary nebula kinematics in M31.

J/ApJ/616/1072/  catalog.dat  (380Kb)

Schmidt M. Astrophys. J., 616, 1072-1077 (2004) The GUSBAD catalog of gamma-ray bursts.

J/ApJ/617/L41/  table1.dat  (10Kb)

Mighell K.J., Roederer I.U. Astrophys. J., 617, L41-L44 (2004) Flickering red giants in the Ursa minor dwarf spheroidal galaxy: detection of low-amplitude variability in faint

red giant branch stars on 10 minute timescales.

J/ApJ/617/184 See II/266/

J/ApJ/617/192/  table1.dat  (29Kb)

Mullis C.R., Henry J.P., Gioia I.M., Bohringer H., Briel U.G., Voges W., Huchra J.P. Astrophys. J., 617, 192-208 (2004) Spatial correlation function of X-ray-selected active galactic nuclei.

J/ApJ/617/240/  table1.dat  (43Kb)

Kobulnicky H.A., Kewley L.J. Astrophys. J., 617, 240-261 (2004) Metallicities of 0.3 z 1.0 galaxies in the GOODS-North field.

J/ApJ/617/262/  table1.dat  (16Kb)

Sivakoff G.R., Sarazin C.L., Carlin J.L. Astrophys. J., 617, 262-280 (2004) Chandra observations of diffuse gas and luminous X-ray sources around the X-ray-bright elliptical galaxy NGC 1600.

J/ApJ/617/718/  refs.dat table2.dat table4.dat  (27Kb)

Prochaska J.X., Chen H.-W., Howk J.C., Weiner B.J., Mulchaey J. Astrophys. J., 617, 718-745 (2004) Probing the intergalactic medium-galaxy connection toward PKS 0405-123. I. Ultraviolet spectroscopy and metal-line systems.

J/ApJ/617/1091/  table3.dat table4.dat  (18Kb)

Simmerer J., Sneden C., Cowan J.J., Collier J., Woolf V.M., Lawler J.E. Astrophys. J., 617, 1091-1114 (2004) The rise of the s-process in the Galaxy.

J/ApJ/618/L83/  refs.dat table1.dat  (14Kb)

Zhou X.-L., Wang J.-M. Astrophys. J., 618, L83-L86 (2005) Narrow iron Kαlines in active galactic nuclei: evolving populations.

J/ApJ/618/123/  table4.dat  (31Kb)

Silverman J.D., Green P.J., Barkhouse W.A., Kim D.-W., Aldcroft T.L., Cameron R.A., Wilkes B.J., Mossman A., Ghosh H., Tananbaum H., Smith M.G., Smith R.C., Smith P.S., Foltz C., Wik D., Jannuzi B.T. Astrophys. J., 618, 123-138 (2005) Hard X-ray-emitting active galactic nuclei selected by the Chandra Multiwavelength Project.

J/ApJ/619/L95/  table1.dat  (9Kb)

Xu C.K., Iglesias-Paramo J., Burgarella D., Rich R.M., Neff S.G., Lauger S., Barlow T.A., Bianchi L., Byun Y.-I., Forster K., Friedman P.G., Heckman T.M., Jelinsky P.N., Lee Y.-W., Madore B.F., Malina R.F., Martin D.C., Milliard B., Morrissey P., Schiminovich D., Siegmund O.H.W., Small T., Szalay A.S., Welsh B.Y., Wyder T.K. Astrophys. J., 619, L95-L98 (2005) Ultraviolet emission and star formation in Stephan’s Quintet.

J/ApJ/619/193/  table6.dat  (60Kb)

McIntosh D.H., Zabludoff A.I., Rix H.-W., Caldwell N. Astrophys. J., 619, 193-217 (2005) Testing the universality of the (U-V) color-magnitude relations for nearby clusters of galaxies.

J/ApJ/619/270/  table3.dat table4.dat table5-9.dat table12.dat  (71Kb)

Melo V.P., Munoz-Tunon C., Maiz-Apellaniz J., Tenorio-Tagle G. Astrophys. J., 619, 270-290 (2005) Young super star clusters in the starburst of M82: the catalog.

J/ApJ/619/755/  table6.dat table7.dat table8.dat  (678Kb)

Dopita M.A., Groves B.A., Fischera J., Sutherland R.S., Tuffs R.J., Popescu C.C., Kewley L.J., Reuland M., Leitherer C. Astrophys. J., 619, 755-778 (2005) Modeling the pan-spectral energy distribution of starburst galaxies. I. The role of ISM pressure and the molecular cloud dissipation timescale.

J/ApJ/620/961/  table1.dat table2.dat table4.dat table5.dat  (31Kb)

Millan-Gabet R., Pedretti E., Monnier J.D., Schloerb F.P., Traub W.A., Carleton N.P., Lacasse M.G., Segransan D. Astrophys. J., 620, 961-969 (2005)

Diameters of Mira stars measured simultaneously in the J, H, and K’ near-infrared bands.

J/ApJ/620/1010/  refs.dat table1.dat  (34Kb)

Rieke G.H., Su K.Y.L., Stansberry J.A., Trilling D., Bryden G., Muzerolle J., White B., Gorlova N., Young E.T., Beichman C.A., Stapelfeldt K.R., Hines D.C. Astrophys. J., 620, 1010-1026 (2005) Decay of planetary debris disks.

J/ApJ/621/53/  table2.dat  (11Kb)

Broadhurst T., Benitez N., Coe D., Sharon K., Zekser K., White R., Ford H., Bouwens R., Blakeslee J., Clampin M., Cross N., Franx M., Frye B., Hartig G., Illingworth G., Infante L., Menanteau F., Meurer G., Postman M., Ardila D.R., Bartko F., Brown R.A., Burrows C.J., Cheng E.S., Feldman P.D., Golimowski D.A., Goto T., Gronwall C., Herranz D., Holden B., Homeier N., Krist J.E., Lesser M.P., Martel A.R., Miley G.K., Rosati P., Sirianni M., Sparks W.B., Steindling S., Tran H.D., Tsvetanov Z.I., Zheng W. Astrophys. J., 621, 53-88 (2005) Strong-lensing analysis of A1689 from deep Advanced Camera images.

J/ApJ/621/104/  table4.dat  (43Kb)

Treister E., Castander F.J., Maccarone T.J., Gawiser E., Coppi P.S., Urry C.M., Maza J., Herrera D., Gonzalez V., Montoya C., Pineda P. Astrophys. J., 621, 104-122 (2005) The Calan-Yale Deep Extragalactic Research (CYDER) survey: optical properties and deep spectroscopy of serendipitous X-ray sources.

J/ApJ/621/673/  table2.dat table3.dat  (23Kb)

Thomas D., Maraston C., Bender R., De Oliveira C.M. Astrophys. J., 621, 673-694 (2005) The epochs of early-type galaxy formation as a function of environment.

J/ApJ/622/129/  objects.dat table1.dat  (310Kb)

Sergeev S.G., Doroshenko V.T., Golubinskiy Y.V., Merkulova N.I., Sergeeva E.A. Astrophys. J., 622, 129-135 (2005) Lag-luminosity relationship for interband lags between variations in B, V, R, and I bands in active galactic nuclei.

J/ApJ/622/187/  table2.dat  (9Kb)

Mahdavi A., Finoguenov A., Bohringer H., Geller M.J., Henry J.P. Astrophys. J., 622, 187-204 (2005) XMM-newton and Gemini observations of eight RASSCALS galaxy groups.

J/ApJ/622/217/  refs.dat table1.dat table2.dat table4.dat  (38Kb)

Whiting A.B. Astrophys. J., 622, 217-234 (2005) Light and motion in the local volume.

J/ApJ/622/294/  stars.dat table2.dat  (21Kb)

Stanghellini L., Shaw R.A., Gilmore D. Astrophys. J., 622, 294-318 (2005) Space telescope imaging spectrograph ultraviolet spectra of Large Magellanic Cloud planetary nebulae: a study of carbon abundances and stellar evolution.

J/ApJ/622/772/  table2.dat  (12Kb)

Chapman S.C., Blain A.W., Smail I., Ivison R.J. Astrophys. J., 622, 772-796 (2005) A redshift survey of the submillimeter galaxy population.

J/ApJ/622/1052/  table1.dat  (25Kb)

McSwain M.V., Gies D.R. Astrophys. J., 622, 1052-1057 (2005) A photometric method to search for Be stars in open clusters.

J/ApJ/622/1102/  table1.dat table2.dat table3.dat  (32Kb)

Fischer D.A., Valenti J. Astrophys. J., 622, 1102-1117 (2005) The planet-metallicity correlation.

J/ApJ/623/181/  table2.dat  (1.4 Mb)

Siegel M.H., Majewski S.R., Gallart C., Sohn S.T., Kunkel W.E., Braun R. Astrophys. J., 623, 181-195 (2005) A search for stellar populations in high-velocity clouds.

J/ApJ/623/897/  refs.dat table1.dat table2.dat  (25Kb)

Larson K.A., Whittet D.C.B. Astrophys. J., 623, 897-910 (2005) Reddening and the extinction law at high galactic latitude.

J/ApJ/624/L5/  table2.dat  (6Kb)

Bouwens R.J., Illingworth G.D., Thompson R.I., Franx M. Astrophys. J., 624, L5-L8 (2005) Constraints on z 10 galaxies from the deepest Hubble Space Telescope NICMOS fields.

J/ApJ/624/135/  table1.dat  (47Kb)

Afonso J., Georgakakis A., Almeida C., Hopkins A.M., Cram L.E., Mobasher B., Sullivan M. Astrophys. J., 624, 135-154 (2005) The Phoenix Deep Survey: spectroscopic catalog.

J/ApJ/625/6/  table3.dat  (10Kb)

Gerke B.F., Newman J.A., Davis M., Marinoni C., Yan R., Coil A.L., Conroy C., Cooper M.C., Faber S.M., Finkbeiner D.P., Guhathakurta P., Kaiser N., Koo D.C., Phillips A.C., Weiner B.J., Willmer C.N.A. Astrophys. J., 625, 6-22 (2005) The DEEP2 Galaxy Redshift Survey: first results on galaxy groups.

J/ApJ/625/78/  table1.dat  (14Kb)

Hao C.N., Xia X.Y., Mao S., Wu H., Deng Z.G. Astrophys. J., 625, 78-88 (2005) The physical connections among infrared QSOs, Palomar-Green QSOs, and narrow-line Seyfert 1 galaxies.

J/ApJ/625/156/  table1.dat  (62Kb)

Dieball A., Knigge C., Zurek D.R., Shara M.M., Long K.S. Astrophys. J., 625, 156-166 (2005) Far-ultraviolet observations of the globular cluster NGC 2808 revisited: blue stragglers, white dwarfs, and cataclysmic variables.

J/ApJ/625/633/  table4.dat table5.dat  (21Kb)

Dai X., Kochanek C.S. Astrophys. J., 625, 633-642 (2005) Chandra observations of the gravitational lenses B1600+434 and B1608+656.

J/ApJ/625/796/  table2.dat table3.dat table4.dat table5.dat table7.dat  (164Kb)

Heinke C.O., Grindlay J.E., Edmonds P.D., Cohn H.N., Lugger P.M., Camilo F., Bogdanov S., Freire P.C. Astrophys. J., 625, 796-824 (2005) A deep Chandra survey of the globular cluster 47 Tucanae: catalog of point sources.

J/ApJ/625/891/  table2.dat table3.dat  (20Kb)

Reid M.A., Wilson C.D. Astrophys. J., 625, 891-905 (2005) High-mass star formation. I. The mass distribution of submillimeter clumps in NGC 7538.

J/ApJ/626/L15/  table1.dat table2.dat  (12Kb)

Hunt L.K., Maiolino R. Astrophys. J., 626, L15-L18 (2005) Improved templates for photometric redshifts of submillimeter sources.

J/ApJ/626/L5/  table1.dat  (6Kb)

Stanek K.Z., Garnavich P.M., Nutzman P.A., Hartman J.D., Garg A., Adelberger K., Berlind P., Bonanos A.Z., Calkins M.L., Challis P., Gaudi B.S., Holman M.J., Kirshner R.P., Mcleod B.A., Osip D., Pimenova T., Reiprich T.H., Romanishin W., Spahr T., Tegler S.C., Zhao X. Astrophys. J., 626, L5-L9 (2005) Deep photometry of GRB 041006 afterglow: hypernova bump at redshift z = 0.716.

J/ApJ/626/95/  table1.dat table2.dat  (61Kb)

Sowards-Emmerd D., Romani R.W., Michelson P.F., Healey S.E., Nolan P.L. Astrophys. J., 626, 95-103 (2005) A northern survey of gamma-ray blazar candidates.

J/ApJ/626/195/  table1.dat table2.dat  (19Kb)

Gibson S.J., Taylor A.R., Higgs L.A., Brunt C.M., Dewdney P.E. Astrophys. J., 626, 195-213 (2005) A self-absorption census of cold H I clouds in the Canadian Galactic Plane Survey.

J/ApJ/626/272/  table2.dat  (15Kb)

Pravdo S.H., Tsuboi Y.

Astrophys. J., 626, 272-282 (2005) X-rays from Cepheus A East and West.

J/ApJ/626/446/  table2.dat table3.dat table4.dat  (27Kb)

Ramirez I., Melendez J. Astrophys. J., 626, 446-464 (2005) The effective temperature scale of FGK stars. I. Determination of temperatures and angular diameters with the infrared flux method.

J/ApJ/626/887/  table2.dat  (10Kb)

Kavars D.W., Dickey J.M., McClure-Griffiths N.M., Gaensler B.M., Green A.J. Astrophys. J., 626, 887-899 (2005) ”Missing link” clouds in the Southern Galactic Plane Survey.

J/ApJ/627/32/  refs.dat table3.dat table4.dat  (39Kb)

Sand D.J., Treu T., Ellis R.S., Smith G.P. Astrophys. J., 627, 32-52 (2005) A systematic search for gravitationally lensed arcs in the Hubble Space Telescope WFPC2 archive.

J/ApJ/627/579/  table2.dat table3.dat table4.dat table5.dat table11.dat table12.dat  (63Kb)

Riess A.G., Li W., Stetson P.B., Filippenko A.V., Jha S., Kirshner R.P., Challis P.M., Garnavich P.M., Chornock R. Astrophys. J., 627, 579-607 (2005) Cepheid calibrations from the Hubble Space Telescope of the luminosity of two recent type Ia supernovae and a redetermination of the Hubble constant.

J/ApJ/627/621/  table1.dat  (11Kb)

Sawicki M., Stevenson M., Barrientos L.F., Gladman B., Mallen-Ornelas G., Van Den Bergh S. Astrophys. J., 627, 621-631 (2005) Properties of extremely red objects in an overdense region.

J/ApJ/628/411/  table4.dat  (4.8 Mb)

Devor J. Astrophys. J., 628, 411-425 (2005) Solutions for 10,000 eclipsing binaries in the bulge fields of OGLE II using DEBiL.

J/ApJ/628/729/  cfht.dat kpno.dat  (2.1 Mb)

Pollard D.L., Sandquist E.L., Hargis J.R., Bolte M. Astrophys. J., 628, 729-737 (2005) The unusual luminosity function of the globular cluster M10.

J/ApJ/628/789/  table1.dat table2.dat table3.dat  (11Kb)

De Gregorio-Monsalvo I., Chandler C.J., Gomez J.F., Kuiper T.B.H., Torrelles J.M., Anglada G. Astrophys. J., 628, 789-799 (2005) High-resolution molecular line observations of the environment of the class 0 source B1-IRS.

J/ApJ/629/288/  table2.dat  (12Kb)

Chen Y., Yao Y., Yang J., Zeng Q., Sato S. Astrophys. J., 629, 288-298 (2005) Near-infrared imaging of the star formation region AFGL 5142.

J/ApJ/629/507/  table1.dat table2.dat table3.dat  (26Kb)

Abt H.A. Astrophys. J., 629, 507-511 (2005) Observed orbital eccentricities.

J/ApJ/629/1055/  refs.dat table1.dat table2.dat table3.dat  (25Kb)

Yakut K., Eggleton P.P. Astrophys. J., 629, 1055-1074 (2005) Evolution of close binary systems.

J/ApJ/630/206/  table3.dat table4.dat table5.dat  (32Kb)

Finn R.A., Zaritsky D., Mccarthy D.W.JR, Poggianti B., Rudnick G., Halliday C., Milvang-jensen B., Pello R., Simard L. Astrophys. J., 630, 206-227 (2005) Hα-derived star formation rates for three z= 0.75 EDisCS galaxy clusters.

J/ApJ/630/381/  table2.dat table3.dat table4.dat table5.dat  (62Kb)

Barsony M., Ressler M.E., Marsh K.A. Astrophys. J., 630, 381-399 (2005) A mid-infrared imaging survey of embedded young stellar objects in the ρOphiuchi cloud core.

J/ApJ/630/784/  table1.dat table4.dat  (102Kb)

Taylor V.A., Jansen R.A., Windhorst R.A., Odewahn S.C., Hibbard J.E. Astrophys. J., 630, 784-803 (2005) UBVR and Hubble Space Telescope mid-ultraviolet and near-infrared surface photometry and radial color gradients of late-type, irregular, and peculiar galaxies.

J/ApJ/631/85/  tablea.dat tableb.dat tablec.dat  (23Kb)

Elmegreen D.M., Elmegreen B.G., Rubin D.S., Schaffer M.A. Astrophys. J., 631, 85-100 (2005) Galaxy morphologies in the Hubble Ultra Deep Field: dominance of linear structures at the detection limit.

J/ApJ/631/1100/  refs.dat table1.dat table2.dat  (15Kb)

Gianninas A., Bergeron P., Fontaine G. Astrophys. J., 631, 1100-1112 (2005) Toward an empirical determination of the ZZ Ceti instability strip.

J/ApJ/632/L131/  table1.dat  (11Kb)

Schuler S.C., Kim J.H., Tinker M.C.Jr, King J.R., Hatzes A.P., Guenther E.W. Astrophys. J., 632, L131-L134 (2005) High-resolution spectroscopy of the planetary host HD 13189: highly evolved and metal-poor.

J/ApJ/632/266/  table2.dat  (656Kb)

Layden A.C., Sarajedini A., von Hippel T., Cool A.M. Astrophys. J., 632, 266-276 (2005) Deep photometry of the globular cluster M5: distance estimates from white dwarf and main-sequence stars.

J/ApJ/632/799/  table1.dat table2.dat  (10Kb)

Peterson B.M., Bentz M.C., Desroches L.-B., Filippenko A.V., Ho L.C., Kaspi S., Laor A., Maoz D., Moran E.C., Pogge R.W., Quillen A.C. Astrophys. J., 632, 799-808 (2005) Multiwavelength monitoring of the dwarf Seyfert 1 galaxy NGC 4395. I. A reverberation-based measurement of the black hole mass.

J/ApJ/633/L77/  table1.dat  (8Kb)

Markwardt C.B., Tueller J., Skinner G.K., Gehrels N., Barthelmy S.D., Mushotzky R.F. Astrophys. J., 633, L77-L80 (2005) The Swift/BAT high-latitude survey: first results.

J/ApJ/633/174/  table1.dat  (36Kb)

Treu T., Ellis R.S., Liao T.X., Van Dokkum P.G., Tozzi P., Coil A., Newman J., Cooper M.C., Davis M. Astrophys. J., 633, 174-197 (2005) The assembly history of field spheroidals: evolution of mass-to-light ratios and signatures of recent star formation.

J/ApJ/633/630/  table1.dat  (7Kb)

Willott C.J., Delfosse X., Forveille T., Delorme P., Gwyn S.D.J. Astrophys. J., 633, 630-637 (2005) First results from the Canada-France high-z quasar survey: constraints on the z=6 quasar luminosity function and the quasar contribution to reionization.

J/ApJ/633/638/  table2.dat  (24Kb)

Wilhite B.C., Vanden Berk D.E., Kron R.G., Schneider D.P., Pereyra N., Brunner R.J., Richards G.T., Brinkmann J.V. Astrophys. J., 633, 638-648 (2005) Spectral variability of quasars in the Sloan Digital Sky Survey. I. Wavelength dependence.

J/ApJ/633/844/  table1.dat  (12Kb)

Pizagno J., Prada F., Weinberg D.H., Rix H.-W., Harbeck D., Grebel E.K., Bell E.F., Brinkmann J., Holtzman J., West A. Astrophys. J., 633, 844-856 (2005) Dark matter and stellar mass in the luminous regions of disk galaxies.

J/ApJ/633/871/  table1.dat  (16Kb)

Calzetti D., Kennicutt R.C.Jr, Bianchi L., Thilker D.A., Dale D.A., Engelbracht C.W., Leitherer C., Meyer M.J., Sosey M.L., Mutchler M., Regan M.W., Thornley M.D., Armus L., Bendo G.J., Boissier S., Boselli A., Draine B.T., Gordon K.D., Helou G., Hollenbach D.J., Kewley L., Madore B.F., Martin D.C., Murphy E.J., Rieke G.H., Rieke M.J., Roussel H., Sheth K., Smith J.D., Walter F., White B.A., Yi S., Scoville N.Z., Polletta M., Lindler D. Astrophys. J., 633, 871-893 (2005) Star formation in NGC 5194 (M51a): the panchromatic view from GALEX to Spitzer.

J/ApJ/634/L1/  table1.dat  (17Kb)

Kasliwal M.M., Charmandaris V., Weedman D., Houck J.R., Le Floc’h E., Higdon S.J.U., Armus L., Teplitz H.I. Astrophys. J., 634, L1-L4 (2005) Identifying silicate-absorbed ULIRGs at z 1-2 in the Bootes field using the Spitzer IRS.

J/ApJ/634/L89/  table1.dat  (6Kb)

Gelfand J.D., Lyubarsky Y.E., Eichler D., Gaensler B.M., Taylor G.B., Granot J., Newton-McGee K.J., Ramirez-Ruiz E., Kouveliotou C., Wijers R.A.M.J. Astrophys. J., 634, L89-L92 (2005) A rebrightening of the radio nebula associated with the 2004 December 27 giant flare from SGR 1806-20.

J/ApJ/634/128/  table1.dat  (11Kb)

Teplitz H.I., Charmandaris V., Chary R., Colbert J.W., Armus L., Weedman D. Astrophys. J., 634, 128-136 (2005) 16μm imaging around the Hubble Deep Field-North with the Spitzer IRS.

J/ApJ/634/272/  table2.dat  (8Kb)

Giordano L., Cortese L., Trinchieri G., Wolter A., Colpi M., Gavazzi G., Mayer L. Astrophys. J., 634, 272-280 (2005) The trail of discrete X-ray sources in the early-type galaxy NGC 4261: anisotropy in the globular cluster distribution?

J/ApJ/634/451/  table1.dat  (14Kb)

Megier A., Strobel A., Bondar A., Musaev F.A., Han I., Krelowski J., Galazutdinov G.A. Astrophys. J., 634, 451-458 (2005) Interstellar Ca II line intensities and the distances of the OB stars.

J/ApJ/634/625/  table1.dat table5.dat  (22Kb)

Rivera E.J., Lissauer J.J., Butler R.P., Marcy G.W., Vogt S.S., Fischer D.A., Brown T.M., Laughlin G., Henry G.W. Astrophys. J. 634, 625 (2005) A 7.5 M(Earth) planet orbiting the nearby star, GJ 876.

J/ApJ/634/861/  table4.dat  (8Kb)

Yamada T., Kodama T., Akiyama M., Furusawa H., Iwata I., Kajisawa M., Iye M., Ouchi M., Sekiguchi K., Shimasaku K., Simpson C., Tanaka I., Yoshida M. Astrophys. J., 634, 861-878 (2005) The number density of old passively evolving galaxies at z=1 in the Subaru/XMM-Newton Deep Survey Field.

J/ApJ/634/977/  table1.dat table2.dat table3.dat  (85Kb)

Moran S.M., Ellis R.S., Treu T., Smail I., Dressler A., Coil A.L., Smith G.P. Astrophys. J., 634, 977-1001 (2005) A Wide-Field Hubble Space Telescope survey of the cluster Cl 0024+16 at z=0.4. III. Spectroscopic signatures of environmental evolution in early-type galaxies.

J/ApJ/634/1319/  table2.dat table4.dat  (25Kb)

Franchini M., Morossi C., Di Marcantonio P., Castelli F., Malagnini M.L., Chavez M. Astrophys. J., 634, 1319-1335 (2005) Synthetic Lick indices and detection of α-enhanced stars. III. F, G, and K stars with [Fe/H] 0.00.

J/ApJ/635/123/  table1.dat table3.dat table4.dat table5.dat  (438Kb)

Prochaska J.X., Herbert-fort S., Wolfe A.M. Astrophys. J., 635, 123-142 (2005) The SDSS damped Lyαsurvey: Data Release 3.

J/ApJ/635/214/  table1.dat table2.dat  (2.5 Mb)

Ebisawa K., Tsujimoto M., Paizis A., Hamaguchi K., Bamba A., Cutri R., Kaneda H., Maeda Y., Sato G., Senda A., Ueno M., Yamauchi S., Beckmann V., Courvoisier T.J.-L., Dubath P., Nishihara E. Astrophys. J., 635, 214-242 (2005) Chandra deep X-ray observation of a typical galactic plane region and near-infrared identification.

J/ApJ/635/243/  table1.dat  (26Kb)

Ferreras I., Lisker T., Carollo C.M., Lilly S.J., Mobasher B. Astrophys. J., 635, 243-259 (2005) Evolution of field early-type galaxies: the view from GOODS CDFS.

J/ApJ/635/280/  table1.dat table2.dat table3.dat  (17Kb)

Brandl B.R., Clark D.M., Eikenberry S.S., Wilson J.C., Henderson C.P., Barry D.J., Houck J.R., Carson J.C., Hayward T.L. Astrophys. J., 635, 280-289 (2005)

Deep near-infrared imaging and photometry of the Antennae galaxies with WIRC.

J/ApJ/635/290/  table5.dat  (25Kb)

Teodorescu A.M., Mendez R.H., Saglia R.P., Riffeser A., Kudritzki R.-P., Gerhard O.E., Kleyna J. Astrophys. J., 635, 290-304 (2005) Planetary nebulae and stellar kinematics in the flattened elliptical galaxy NGC 1344.

J/ApJ/635/560/  table1.dat  (10Kb)

Mercer E.P., Clemens D.P., Meade M.R., Babler B.L., Indebetouw R., Whitney B.A., Watson C., Wolfire M.G., Wolff M.J., Bania T.M., Benjamin R.A., Cohen M., Dickey J.M., Jackson J.M., Kobulnicky H.A., Mathis J.S., Stauffer J.R., Stolovy S.R., Uzpen B., Churchwell E.B. Astrophys. J., 635, 560-569 (2005) New star clusters discovered in the GLIMPSE survey.

J/ApJ/635/950/  notes.dat table2.dat table3.dat  (14Kb)

Henkel B., Partridge R.B. Astrophys. J., 635, 950-958 (2005) Completing the counts of radio sources at 8.5GHz.

J/ApJ/635/1053/  table1.dat  (10Kb)

David L.P., Jones C., Forman W., Murray S.S. Astrophys. J., 635, 1053-1061 (2005) The diffuse emission and a variable ultraluminous X-ray point source in the elliptical galaxy NGC 3379.

J/ApJ/636/200/  table1.dat  (13Kb)

Randall S.W., Sarazin C.L., Irwin J.A. Astrophys. J., 636, 200-213 (2006) XMM-Newton observation of diffuse gas and low-mass X-ray binaries in the elliptical galaxy NGC 4649 (M60).

J/ApJ/636/214/  table1.dat table2.dat table3.dat table4.dat table5.dat table6.dat  (48Kb)

van Zee L., Haynes M.P. Astrophys. J., 636, 214-239 (2006) Oxygen and nitrogen in isolated dwarf irregular galaxies.

J/ApJ/636/240/  tablea1.dat tablea2.dat tablea3.dat tablea4/* tablea5.dat  (2.3 Mb)

Sumi T., Wozniak P.R., Udalski A., Szymanski M., Kubiak M., Pietrzynski G., Soszynski I., Zebrun K., Szewczyk O., Wyrzykowski L., Paczynski B. Astrophys. J., 636, 240-260 (2006) Microlensing optical depth toward the Galactic Bulge using bright sources from OGLE-II.

J/ApJ/636/610/  refs.dat table1.dat  (33Kb)

Rao S.M., Turnshek D.A., Nestor D.B. Astrophys. J., 636, 610-630 (2006) Damped Lyαsystems at z 1.65: the expanded Sloan Digital Sky Survey Hubble Space Telescope sample.

J/ApJ/636/631/  table1.dat table4.dat table5.dat  (46Kb)

Williger G.M., Heap S.R., Weymann R.J., Dave R., Ellingson E., Carswell R.F., Tripp T.M., Jenkins E.B. Astrophys. J., 636, 631-653 (2006) The low-redshift Lyαforest toward PKS 0405-123.

J/ApJ/636/721/  table1.dat table3.dat  (26Kb)

Brownstein J.R., Moffat J.W. Astrophys. J., 636, 721-741 (2006) Galaxy rotation curves without nonbaryonic dark matter.

J/ApJ/636/932/  refs.dat table1.dat table3.dat table4.dat  (49Kb)

McCabe C., Ghez A.M., Prato L., Duchene G., Fisher R.S., Telesco C. Astrophys. J., 636, 932-951 (2006) Investigating disk evolution: a high spatial resolution mid-infrared survey of T Tauri stars.

J/ApJ/636/1002/  table1.dat table3.dat  (118Kb)

Sokoloski J.L., Kenyon S.J., Espey B.R., Keyes C.D., Mccandliss S.R., Kong A.K.H., Aufdenberg J.P., Filippenko A.V., Li W., Brocksopp C., Kaiser C.R., Charles P.A., Rupen M.P., Stone R.P.S. Astrophys. J., 636, 1002-1019 (2006) A “Combination nova” outburst in Z Andromedae: nuclear shell burning triggered by a disk instability.

J/ApJ/636/1020/  table1.dat table2.dat table3.dat table4.dat  (64Kb)

Muterspaugh M.W., Lane B.F., Konacki M., Wiktorowicz S., Burke B.F., Colavita M.M., Kulkarni S.R., Shao M. Astrophys. J., 636, 1020-1032 (2006) PHASES differential astrometry and iodine cell radial velocities of the κPegasi triple star system.

J/ApJ/636/1098/  refs.dat table1.dat table2.dat  (25Kb)

Bryden G., Beichman C.A., Trilling D.E., Rieke G.H., Holmes E.K., Lawler S.M., Stapelfeldt K.R., Werner M.W., Gautier T.N., Blaylock M., Gordon K.D., Stansberry J.A., Su K.Y.L. Astrophys. J., 636, 1098-1113 (2006) Frequency of debris disks around solar-type stars: first results from a Spitzer MIPS survey.

J/ApJ/637/1067/  refs.dat table13.dat table14.dat  (27Kb)

Burgasser A.J., Geballe T.R., Leggett S.K., Kirkpatrick J.D., Golimowski D.A. Astrophys. J., 637, 1067-1093 (2006) A unified near-infrared spectral classification scheme for T dwarfs.

J/ApJ/637/1102/  table2.dat  (35Kb)

Robinson S.E., Strader J., Ammons S.M., Laughlin G., Fischer D. Astrophys. J., 637, 1102-1112 (2006) The N2K consortium. V. Identifying very metal-rich stars with low-resolution spectra: finding planet-search targets.

J/ApJ/638/897/  table2.dat table3.dat table4.dat table5.dat table6.dat  (64Kb)

Sicilia-Aguilar A., Hartmann L., Calvet N., Megeath S.T., Muzerolle J., Allen L., D’Alessio P., Merin B., Stauffer J., Young E., Lada C. Astrophys. J., 638, 897-919 (2006) Disk evolution in Cep OB2: results from the Spitzer Space Telescope.

J/ApJ/639/37/  table1.dat table2.dat  (17Kb)

Koulouridis E., Plionis M., Chavushyan V., Dultzin-Hacyan D., Krongold Y., Goudis C. Astrophys. J., 639, 37-45 (2006) Local and large-scale environment of Seyfert galaxies.

J/ApJ/639/227/  catalog.dat  (1.7 Mb)

Simon R., Jackson J.M., Rathborne J.M., Chambers E.T. Astrophys. J., 639, 227-236 (2006) A catalog of Midcourse Space Experiment infrared dark cloud candidates.

J/ApJ/639/259/  table1.dat table2.dat  (14Kb)

Johnstone D., Matthews H., Mitchell G.F. Astrophys. J., 639, 259-274 (2006) Large area mapping at 850 μm. IV. Analysis of the clump distribution in the Orion B South molecular cloud.

J/ApJ/639/724/  table2.dat  (11Kb)

Straughn A.N., Cohen S.H., Ryan R.E., Hathi N.P., Windhorst R.A., Jansen R.A. Astrophys. J., 639, 724-730 (2006) Tracing galaxy assembly: tadpole galaxies in the Hubble Ultra Deep Field.

J/ApJ/639/766/  table1.dat  (468Kb)

Prochter G.E., Prochaska J.X., Burles S.M. Astrophys. J., 639, 766-780 (2006) On the incidence and kinematics of strong Mg II absorbers.

J/ApJ/639/1069/  stars.dat table2.dat  (8Kb)

Hillwig T.C., Gies D.R., Bagnuolo W.G.JR, Huang W., Mcswain M.V., Wingert D.W. Astrophys. J., 639, 1069-1080 (2006) Binary and multiple O-type stars in the Cassiopeia OB6 association.

J/ApJ/640/L43/  table1.dat  (9Kb)

Sollima A., Borissova J., Catelan M., Smith H.A., Minniti D., Cacciari C., Ferraro F.R. Astrophys. J., 640, L43-L46 (2006) New metallicities of RR Lyrae stars in ωCentauri: evidence for a non-He-enhanced metal-intermediate population.

J/ApJ/640/603/  table1.dat  (31Kb)

Treister E., Urry C.M., Van Duyne J., Dickinson M., Chary R.-R., Alexander D.M., Bauer F., Natarajan P., Lira P., Grogin N.A. Astrophys. J., 640, 603-611 (2006) Spitzer number counts of active galactic nuclei in the GOODS fields.

J/ApJ/640/982/  table2.dat table3.dat table4.dat  (35Kb)

Chen X., Shen Z.-Q., Imai H., Kamohara R. Astrophys. J., 640, 982-994 (2006) Inward motions of the compact SiO masers around VX Sagittarii.

J/ApJ/640/1018/  table1.dat table3.dat table4.dat table5.dat table6.dat table7.dat  (219Kb)

Torres G., Lacy C.H., Marschall L.A., Sheets H.A., Mader J.A. Astrophys. J., 640, 1018-1038 (2006) The eclipsing binary V1061 Cygni: confronting stellar evolution models for active and inactive solar-type stars.

J/ApJ/640/1051/  table1.dat  (18Kb)

Grether D., Lineweaver C.H. Astrophys. J., 640, 1051-1062 (2006) How dry is the brown dwarf desert? Quantifying the relative number of planets, brown dwarfs, and stellar companions around nearby Sun-like stars.

J/ApJ/641/L13/  table1.dat  (23Kb)

Gorosabel J., Castro-Tirado A.J., Ramirez-Ruiz E., Granot J., Caon N., Cairos L.M., Rubio-Herrera E., Guziy S., De Ugarte Postigo A., Jelinek M. Astrophys. J., 641, L13-L16 (2006) Revealing the jet structure of GRB 030329 with high-resolution multicolor photometry.

J/ApJ/641/L21/  table1.dat  (13Kb)

Greene J.E., Ho L.C. Astrophys. J., 641, L21-L24 (2006) The M_{BH}-σ_{∗} relation in local active galaxies.

J/ApJ/641/78/  table1.dat table2.dat table3.dat  (22Kb)

Wilhite B.C., Vanden Berk D.E., Brunner R.J., Brinkmann J.V. Astrophys. J., 641, 78-89 (2006) Spectral variability of quasars in the Sloan Digital Sky Survey. II. The C IV line.

J/ApJ/641/140/  table2.dat  (1.3 Mb)

Brand K., Brown M.J.I., Dey A., Jannuzi B.T., Kochanek C.S., Kenter A.T., Fabricant D., Fazio G.G., Forman W.R., Green P.J., Jones C.J., Mcnamara B.R., Murray S.S., Najita J.R., Rieke M., Shields J.C., Vikhlinin A. Astrophys. J., 641, 140-157 (2006) The Chandra XBootes survey. III. Optical and near-infrared counterparts.

J/ApJ/642/126/  table1.dat  (20Kb)

Barmby P., Alonso-herrero A., Donley J.L., Egami E., Fazio G.G., Georgakakis A., Huang J.-S., Laird E.S., Miyazaki S., Nandra K., Park S.Q., Perez-gonzalez P.G., Rieke G.H., Rigby J.R., Willner S.P. Astrophys. J., 642, 126-139 (2006) Mid-infrared properties of X-ray sources in the extended Groth strip.

J/ApJ/643/680/  fits/* pks0405.dat table1.dat  (103.7 Mb)

Prochaska J.X., Weiner B.J., Chen H.-W., Mulchaey J.S. Astrophys. J., 643, 680-691 (2006) Probing the intergalactic medium-galaxy connection toward PKS 0405-123. III. The galaxy survey and correlations with O VI absorbers.

J/ApJ/643/965/  table1.dat table2.dat  (28Kb)

Rho J., Reach W.T., Lefloch B., Fazio G.G. Astrophys. J., 643, 965-977 (2006) Spectacular Spitzer images of the Trifid nebula: protostars in a young, massive-star-forming region.

J/ApJ/643/1011/  table2.dat  (19Kb)

Paumard T., Genzel R., Martins F., Nayakshin S., Beloborodov A.M., Levin Y., Trippe S., Eisenhauer F., Ott T., Gillessen S., Abuter R., Cuadra J., Alexander T., Sternberg A. Astrophys. J., 643, 1011-1035 (2006) The two young star disks in the central parsec of the Galaxy: properties, dynamics, and formation.

J/ApJ/644/30/  table1.dat table2.dat  (33Kb)

Blakeslee J.P., Holden B.P., Franx M., Rosati P., Bouwens R.J., Demarco R., Ford H.C., Homeier N.L., Illingworth G.D., Jee M.J., Mei S., Menanteau F., Meurer G.R., Postman M., Tran K.-V.H. Astrophys. J., 644, 30-53 (2006) Clusters at half Hubble time: galaxy structure and colors in RX J0152.7-1357 and MS 1054-03.

J/ApJ/644/100/  table3.dat  (27Kb)

Prescott M.K.M., Impey C.D., Cool R.J., Scoville N.Z. Astrophys. J., 644, 100-115 (2006) Quasars in the COSMOS field.

J/ApJ/644/829/  table2.dat table456.dat  (12Kb)

Kim D.-W., Barkhouse W.A., Romero-Colmenero E., Green P.J., Kim M., Mossman A., Schlegel E., Silverman J.D., Aldcroft T., Anderson C., Ivezic Z., Kashyap V., Tananbaum H., Wilkes B.J. Astrophys. J., 644, 829-842 (2006) Chandra Multiwavelength Project: normal galaxies at intermediate redshift.

J/ApJ/644/990/  table1.dat table2.dat  (30Kb)

Reid M.A., Wilson C.D. Astrophys. J., 644, 990-1005 (2006) High-mass star formation. II. The mass function of submillimeter clumps in M17.

J/ApJ/644/1006/  table1.dat  (47Kb)

Blagrave K.P.M., Martin P.G., Baldwin J.A. Astrophys. J., 644, 1006-1027 (2006) A photoionized Herbig-Haro object in the Orion nebula.

J/ApJ/645/228/  table3.dat  (11Kb)

Azzaro M., Zentner A.R., Prada F., Klypin A.A. Astrophys. J., 645, 228-239 (2006) The peculiar velocities of satellites of external disk galaxies.

J/ApJ/645/256/  table2.dat  (12Kb)

Isobe N., Makishima K., Tashiro M., Itoh K., Iyomoto N., Takahashi I., Kaneda H. Astrophys. J., 645, 256-263 (2006) The XMM-Newton examination of energetics in the east lobe of the nearby radio galaxy Fornax A (NGC 1316).

J/ApJ/645/488/  refs.dat table1.dat  (23Kb)

Wang X., Wang L., Pain R., Zhou X., Li Z. Astrophys. J., 645, 488-505 (2006) Determination of the Hubble constant, the intrinsic scatter of luminosities of type Ia supernovae, and evidence for nonstandard dust in other galaxies.

J/ApJ/645/613/  table1.dat table2.dat table4.dat table6.dat  (150Kb)

Ivans I.I., Simmerer J., Sneden C., Lawler J.E., Cowan J.J., Gallino R., Bisterzo S. Astrophys. J., 645, 613-633 (2006) Near-ultraviolet observations of HD 221170: new insights into the nature of r-process-rich stars.

J/ApJ/645/676/  table1.dat table2.dat table3.dat  (18Kb)

Luhman K.L. Astrophys. J., 645, 676-687 (2006) The spatial distribution of brown dwarfs in Taurus.

J/ApJ/645/856/  table1.dat  (15Kb)

Wang T.-G., Zhou H.-Y., Wang J.-X., Lu Y.-J., Lu Y. Astrophys. J., 645, 856-860 (2006) Evidence for a population of beamed radio-intermediate quasars.

J/ApJ/645/890/  table1.dat  (12Kb)

Wang R., Wu X.-B., Kong M.-Z. Astrophys. J., 645, 890-899 (2006) The black hole fundamental plane from a uniform sample of radio and X-ray-emitting broad-line AGNs.

J/ApJ/648/7/  table2.dat  (7Kb)

Kashikawa N., Shimasaku K., Malkan M.A., Doi M., Matsuda Y., Ouchi M., Taniguchi Y., Ly C., Nagao T., Iye M., Motohara K., Murayama T., Murozono K., Nariai K., Ohta K., Okamura S., Sasaki T., Shioya Y., Umemura M. Astrophys. J., 648, 7-22 (2006) The end of the reionization epoch probed by Lyαemitters at z = 6.5 in the Subaru Deep Field.

J/ApJ/654/915/  fors1.dat  (4.7 Mb)

Sollima A., Ferraro F.R., Bellazzini M., Origlia M., Straniero O., Pancino E. Astrophys. J. 654, 915-922 (2007) Deep FORS1 observations of the double Main Sequence of omega Centauri.

J/ApJ/661/750/  table1.dat  (23Kb)

Tran K.-V.H., Franx M., Illingworth G.D., Van Dokkum P., Kelson D.D., Blakeslee J.P., Postman M. Astrophys. J., 661, 750-767 (2007) A Keck spectroscopic survey of MS 1054-03 (z = 0.83): forming the red sequence.

J/ApJS/2/75 See II/221A/

J/ApJS/2/123 See II/221A/

J/ApJS/2/195 See V/55/

J/ApJS/2/271 See II/221A/ J/ApJS/2/298 See II/221A/ J/ApJS/3/211 See VII/4A/ J/ApJS/4/1 See II/221A/ J/ApJS/4/23 See II/221A/ J/ApJS/4/73 See II/221A/ J/ApJS/4/257 See VII/20/ J/ApJS/7/1 See VII/7A/ J/ApJS/12/163 See VII/9/ J/ApJS/13/65/ notes.dat table4.dat (84Kb)

Pauliny-Toth I.I.K., Wade C.M., Heeschen D.S.

Astrophys. J. Suppl. Ser. 13, 65 (1966) Positions and flux densities of radio sources.

J/ApJS/14/1 See VII/74A/ J/ApJS/17/371 See III/23B/ J/ApJS/18/47 See II/42/ J/ApJS/19/367 See III/12/ J/ApJS/20/1 See VII/2A/ J/ApJS/22/177 See III/27/ J/ApJS/23/421 See II/43/ J/ApJS/23/453 See II/44/ J/ApJS/27/37 See IX/3/ J/ApJS/32/7 See III/48B/ J/ApJS/33/113 See VII/17/ J/ApJS/33/269 See III/77/ J/ApJS/33/459 See III/180/ J/ApJS/34/95 See VII/171/ J/ApJS/34/115 See II/36/ J/ApJS/34/381 See VII/87A/ J/ApJS/35/197 See VII/171/ J/ApJS/35/203 See VII/171/ J/ApJS/35/209 See VII/28/ J/ApJS/36/53/ table23.dat (70Kb)

Dressel L.L., Condon J.J.

Astrophys. J. Suppl. Ser. 36, 53 (1978)

The Arecibo 2380 MHz survey of bright galaxies. J/ApJS/36/587 See VII/171/ J/ApJS/38/287 See V/47/ J/ApJS/38/351/ refstars.dat table2.dat (20Kb)

Ford H.C., Jacoby G.H.

Astrophys. J. Suppl. Ser. 38, 351 (1978)

Planetary nebulae in local-group galaxies. VIII. A catalog of planetary nebulae in the Andromeda galaxy. J/ApJS/38/357 See VII/18/ J/ApJS/39/195 See II/67/ J/ApJS/40/1 See VI/39/ J/ApJS/40/3 See VII/33/ J/ApJS/40/583 See VIII/3/ J/ApJS/42/565 See VII/174/ J/ApJS/43/57 See VII/27/ J/ApJS/43/501 See II/83/

J/ApJS/45/97/  table2.dat  (38Kb)

Simard-Normandin M., Kronberg P.P, Button S.

Astrophys. J. Suppl. 45, 97 (1981)

The Faraday Rotation Measures of Extragalactic Radio Sources J/ApJS/45/113 See VII/171/ J/ApJS/45/437 See III/106/ J/ApJS/46/117 See II/115A/ J/ApJS/48/51 See II/115A/ J/ApJS/49/27 See II/115A/ J/ApJS/49/183 See VII/50/ J/ApJS/50/241 See VII/75/ J/ApJS/50/451 See II/88/ J/ApJS/51/67/ table1.dat table2.dat table4.dat (323Kb)

Lawrence C.R., Bennett C.L., Garcia-Barreto J.A., Greenfield P.E., Burke B.F.

Astrophys. J. Suppl. Ser. 51, 67 (1983)

5 GHz observations of sources in the Arecibo 611 MHz survey J/ApJS/52/61 See VII/84A/ J/ApJS/52/121 See III/88/ J/ApJS/53/413 See II/95/ J/ApJS/53/791/ table1.dat table2.dat table3.dat table5.dat (25Kb)

Pier J.R.

Astrophys. J. Suppl. Ser. 53, 791 (1983)

AB stars in the southern galactic halo. II. Spectroscopy and radial velocities J/ApJS/55/657 See III/117/ J/ApJS/56/257 See III/92/ J/ApJS/56/393 See II/115A/ J/ApJS/56/403/ refs table1 table2 table3 table4 table5a table5b table5c table6a table6b table6c (541Kb)

KOWALSKI M.P., ULMER M.P., CRUDDACE R.G., WOOD K.S.

Astrophys. J. Suppl. Ser. 56, 403 (1984)

An X-ray survey of clusters of galaxies. IV. A survey of southern clusters and a compilation of upper limits for both Abell and Southern clusters.

J/ApJS/56/507 See VII/80/ J/ApJS/57/261 See VII/64/ J/ApJS/58/561 See VI/55/ J/ApJS/58/711 See VI/55/ J/ApJS/59/1 See VII/72/ J/ApJS/59/33 See VII/102/ J/ApJS/59/95/ table1.dat (85Kb)

Abt H.A.

Astrophys. J. Suppl. Ser., 59, 95-112 (1985)

Visual multiples. VIII. 1000 MK types. J/ApJS/59/197 See III/119/ J/ApJS/59/397 See II/114/ J/ApJS/60/297 See VII/114/ J/ApJS/61/1 See VIII/52B/ J/ApJS/61/305 See II/207/ J/ApJS/61/667/ table1.dat (87Kb)

Norris J.

Astrophys. J. Suppl. Ser. 61, 667 (1986)

Population Studies. II. Kinematics as a Function of Abundance and Galactocentric Position for [Fe/H] = -0.6 J/ApJS/62/501 See III/114/ J/ApJS/62/673 See I/123/ J/ApJS/63/1 See VII/105/

J/ApJS/63/335 See VI/55/

J/ApJS/63/543 See VII/95/

J/ApJS/63/555 See VII/96/

J/ApJS/63/661/ table1.dat table2.dat table3.dat table4.dat table5.dat table6.dat table7.dat table8.dat table9.dat table10.dat (175Kb) Nakagawa M., Kohyama Y., Itoh N. Astrophys. J. Suppl. Ser. 63, 661 (1987) (From AAS CD-ROM series, Volume 9, 1997 Lee E. Brotzman [ADS] 28-Aug-97) Relativistic Free-Free Gaunt Factor of the Dense High-Temperature Stellar Plasma

J/ApJS/63/821/  table1.dat table2.dat  (199Kb)

Scoville N.Z., Yun Min Su, Sanders D.B., Clemens D.P., Waller W.H. Astrophys. J. Suppl. Ser. 63, 821 (1987) Molecular clouds and cloud cores in the inner Galaxy

J/ApJS/64/103 See VI/55/

J/ApJS/64/127 See VII/94/

J/ApJS/65/459 See III/136/

J/ApJS/65/485/  notes.dat table1.dat table2.dat  (112Kb)

Condon J.J. Astrophys. J. Suppl. 65, 485-541 (1987) A 1.49 GHz atlas of spiral galaxies with B_{T} =12 and δ=-45^{∘}.

J/ApJS/65/603 See III/129/

J/ApJS/66/1 See II/115A/

J/ApJS/66/309/  table1  (13Kb)

SANDULEAK N. Astrophys. J. Suppl. Ser. 66, 309 (1988) The Case low-dispersion Northern Sky Survey. VI. A-F stars in a region near the North Galactic Pole

J/ApJS/66/387/  table1.dat  (10Kb)

Sanduleak N., Pesch P. Astrophys. J. Suppl. Ser. 66, 387 (1988) The case low-dispersion northern sky survey. VII. Late-type stars.

J/ApJS/68/1 See II/143A/

J/ApJS/68/91 See VII/109/

J/ApJS/68/715/  table4.dat  (29Kb)

Kassim N.E. Astrophys. J. Suppl. Ser. 68, 715 (1988) The Clark Lake 30.9 MHz Galactic plane survey

J/ApJS/69/1 See VII/117/

J/ApJS/69/99/  table2 table3  (54Kb)

Straw S.M., Hyland A.R., McGregor P.J. Astrophys. J. Suppl. Ser. 69, 99 (1989) The centers of star formation in NGC 6334 and their stellar mass distributions

J/ApJS/69/651 See I/142/

J/ApJS/69/763 See VII/176A/

J/ApJS/70/1 See VII/110A/

J/ApJS/70/329/  calc.for table2  (104Kb)

KNAPP G.R., GUHATHAKURTA P., KIM D.-W., JURA M. Astrophys. J. Suppl. Ser. 70, 329 (1989) Interstellar matter in early-type galaxies. I. IRAS flux densities

J/ApJS/70/687 See VII/213/

J/ApJS/70/833 See VIII/12/

J/ApJS/71/25/  table3 table4  (71Kb)

Brocato E., Buonanno R., Castellani V., Walker A.R. Astrophys. J. Suppl. Ser. 71, 25 (1989)

CCD photometry of the Large Magellanic cloud cluster NGC 1866: observational results

J/ApJS/71/245 See III/150A/

J/ApJS/71/469 See VIII/58/

J/ApJS/71/549/  table1  (28Kb)

PESCH P., SANDULEAK N. Astrophys. J. Suppl. Ser. 71, 549 (1989) The Case low-dispersion Northern Sky Survey. X. A-F stars in a region at intermediate galactic latitude primarily in Lynx and Leo Minor

J/ApJS/71/983/  table2  (39Kb)

VIOTTI R., ROSSI L., CASSATELLA A., ALTAMORE A., BARATTA G.B. Astrophys. J. Suppl. Ser. 71, 983 (1989) The ultraviolet spectrum of eta Carinae

J/ApJS/72/471 See VII/181/

J/ApJS/72/567 See IX/15/

J/ApJS/72/621 See VIII/52B/

J/ApJS/73/359/  iras1490.dat  (63Kb)

Condon J.J., Helou G., Sanders D.B., Soifer B.T. Astrophys. J. Suppl. Ser. 73, 359 (1990) A 1.49 GHz atlas of the IRAS bright galaxy sample.

J/ApJS/73/661/  table1 table2 table3  (57Kb)

Hodge P., Gurwell M., Goldader J.D., Kennicutt R.C. Jr. Astrophys. J. Suppl. Ser. 73, 661 (1990) The H II regions of M 101. I. An atlas of 1264 emission regions

J/ApJS/73/671/  table2  (17Kb)

COUTURE J., HARRIS W.E., ALLWRIGHT J.W.B. Astrophys. J. Suppl. Ser. 73, 671 (1990) BVI photometry of globular clusters in M87

J/ApJS/73/781/  snr ctrt.dat snr indx.dat snr psrc.dat  (36Kb)

Seward F.D. Astrophys. J. Suppl. Ser. 73, 781 (1990) EINSTEIN observations of galactic supernova remnants.

J/ApJS/74/129 See VIII/52B/

J/ApJS/74/181/  table4.dat table8.dat  (145Kb)

Zoonematkermani S., Helfand D.J., Becker R.H., White R.L., Perley R.A. Astrophys. J. Suppl. Ser. 74, 181 (1990) A catalog of small-diameter radio sources in the Galactic plane

J/ApJS/74/275/  notes tables  (87Kb)

Heintz W.D. Astrophys. J. Suppl. Ser. 74, 275 (1990) Observations of double stars and new pairs. XIV.

J/ApJS/74/291/ table1.dat table2.dat table3.dat table4.dat table5.dat table6.dat table7.dat table8.dat table9.dat table10.dat (150Kb) Itoh N., Kojo K., Nakagawa M. Astrophys. J. Suppl. Ser. 74, 291 (1990) (From AAS CD-ROM series, Volume 9, 1997 Lee E. Brotzman [ADS] 28-Aug-97) Relativistic Free-Free Gaunt Factor of the Dense High-Temperature Stellar Plasma. II. Carbon and Oxygen Plasmas

J/ApJS/74/885 See II/115A/

J/ApJS/74/911 See V/68A/

J/ApJS/75/1 See VIII/13/

J/ApJS/75/241/  table1.dat  (150Kb)

Dressler A. Astrophys. J. Suppl. Ser. 75, 241 (1991) The supergalactic plane redshift survey

J/ApJS/75/297 See VII/178/

J/ApJS/75/407/  table4a  (59Kb)

Meatheringham S.J., Dopita M.A. Astrophys. J. Suppl. Ser. 75, 407 (1991) Optical spectroscopy of Magellanic Cloud planetary nebulae. I.

J/ApJS/75/645 See III/157/

J/ApJS/75/801/  table4.dat  (0Kb)

Griffith M., Langston G., Heflin M., Conner S., Burke B. Astrophys. J. Suppl. Ser. 75, 801 (1991) The Fourth MIT-Green Bank 5GHz Survey

J/ApJS/75/1011 See VIII/14/

J/ApJS/76/1 See VII/182/

J/ApJS/76/23 See VII/148/

J/ApJS/76/383 See III/159A/

J/ApJS/76/813 See IX/15/

J/ApJS/76/979 See VI/51/

J/ApJS/76/1001/  table3.dat table4.dat table5.dat  (105Kb)

Preston G.W., Shectman S.A., Beers T.C. Astrophys. J. Suppl. Ser. 76, 1001 (1991) Photoelectric UBV photometry of stars selected in the HK objective-prism survey

J/ApJS/76/1085/  table4  (56Kb)

Meatheringham S.J., Dopita M.A. Astrophys. J. Suppl. Ser. 76, 1085 (1991) Optical spectroscopy of Magellanic Cloud planetary nebulae. II.

J/ApJS/77/203/  refalp refs table1a table1b table2 table3  (355Kb)

JUNKKARINEN V., HEWITT A., BURBIDGE G. Astrophys. J. Suppl. Ser. 77, 203 (1991) (Erratum in 1992, Ap. J. Supplement, 81, 409) A catalog of absorption in the spectra of quasi-stellar objects.

J/ApJS/77/363 See VII/177/

J/ApJS/77/473 See VI/55/

J/ApJS/78/1/ table2.dat table3.dat table4.dat table5.dat table6.dat table7.dat table8.dat (242Kb) Dressler A., Gunn J.E. Astrophys. J. Suppl. Ser. 78, 1 (1992) Spectroscopy of galaxies in distant clusters. IV. A catalog of photometry and spectroscopy for galaxies in seven clusters with 0.35 z 0.55

J/ApJS/78/239/  table4 table5  (118Kb)

WEAVER W.B., JONES G. Astrophys. J. Suppl. Ser. 78, 239 (1992) (From AAS CD-ROM series, Volume 2, 1994) A catalog of CO-added IRAS fluxes of Orion population stars

J/ApJS/79/77 See VIII/10/

J/ApJS/79/255 See VII/141/

J/ApJS/79/331 See VIII/17/

J/ApJS/80/137/  southext.dat  (42Kb)

Jones P.A., McAdam W.B. Astrophys. J. Suppl. Ser. 80, 137 (1992) The structure of southern extragalactic radio sources.

J/ApJS/80/479/  notes.dat table1.dat table2.dat  (58Kb)

Tully R.B., Shaya E.J., Pierce M.J. Astrophys. J. Suppl. Ser. 80, 479 (1992) Nearby galaxy flows modeled by the light distribution: distances, model, and the local velocity anomaly.

J/ApJS/80/501/  table1  (14Kb)

Owen F.N., White R.A., Burns J.O. Astrophys. J. Suppl. Ser. 80, 501 (1992) A 20 centimeter VLA survey of Abell clusters of galaxies. II. Images and optical identifications

J/ApJS/80/531/  gxatlas.dat gxfluxes.dat  (111Kb)

Fabbiano G., Kim D.-W., Trinchieri G. Astrophys. J. Suppl. Ser. 80, 531 (1992)

An X-ray catalog and atlas of galaxies.

J/ApJS/81/5/  notes.dat table1.dat table2.dat table3.dat  (90Kb)

Schneider S.E., Thuan T.X., Mangum J.G., Miller J. Astrophys. J. Suppl. Ser. 81, 5 (1992) Northern dwarf and low surface brightness galaxies. II. The Green Bank neutral hydrogen survey.

J/ApJS/81/83/  table1.dat  (29Kb)

Herbig T., Readhead A.C.S. Astrophys. J. Suppl. 81, 83 (1992) Compendium of Radio Spectra and Luminosities for three complete samples of Radio Sources

J/ApJS/81/163 See VI/55/

J/ApJS/81/221 See VI/54/

J/ApJS/81/413/  table1.dat table2a.dat table2b.dat  (362Kb)

Mathewson D.S., Ford V.L., Buchhorn M. Astrophys. J. Suppl. Ser. 81, 413 (1992) A southern sky survey of the peculiar velocities of 1355 spiral galaxies.

J/ApJS/81/865 See III/166/

J/ApJS/82/117/  table4 table4.tex table5 table5.tex  (628Kb)

Slawson R.W., Hill R.J., Landstreet J.D. Astrophys. J. Suppl. Ser. 82, 117 (1992) A homogeneous catalog of new UBV and H-beta photometry of B and A type stars in and around the Scorpius-Centaurus OB association

J/ApJS/83/29 See II/174/

J/ApJS/83/87/  table4a  (63Kb)

Vassiliadis E., Dopita M.A., Morgan D.H., Bell J.F. Astrophys. J. Suppl. Ser. 83, 87 (1992) Optical spectroscopy of Magellanic Cloud planetary nebulae. III.

J/ApJS/83/351/  table1 table2 table4  (74Kb)

HEINTZ W.D. Astrophys. J. Suppl. Ser. 83, 351 (1992) Observations of double stars and new pairs. XV.

J/ApJS/85/181 See VI/59/

J/ApJS/85/249/  table1.dat  (40Kb)

Hamabe M. Astrophys. J. Suppl. 85, 249 (1993) Surface Photometry of the Hydra I Cluster of Galaxies. I. Photometric Data.

J/ApJS/85/293/  table1 table2 table3  (146Kb)

DE MARCHI G., PARESCE F., FERRARO F.R. Astrophys. J. Suppl. Ser. 85, 293 (1993) The structure of the core of 47 Tucanae from Hubble Space Telescope observations

J/ApJS/85/315/  table7a table7b  (28Kb)

SODERBLOM D.R., STAUFFER J.R., HUDON J.D., JONES B.F. Astrophys. J. Suppl. Ser. 85, 315 (1993) Rotation and Chromospheric emission among F, G and K dwarf stars of the Pleiades

J/ApJS/86/5 See III/167/

J/ApJS/86/453/  table.dat  (25Kb)

MacConnell J.D., Stephenson C. B., Pesch P. Astrophys. J. Suppl. Ser. 86, 453 (1993) The Case Low-Dispersion Northern Sky Survey. XIV. A-F Stars in a Region in Southern Canes Venatici

J/ApJS/86/713/ fig2 fig3 fig4 fig5 fig6 fig7 fig8 fig9 fig10 fig11 fig12 fig13 fig14 fig15 fig16 fig17 fig18 fig19 table2 table3 table4 table5 table6 table7 table8 table9 table10 table11 table12 table13 table14 table15 table16 table17 table18 table19 (17.8 Mb) HUDGINS D.M., SANDFORD S.A., ALLAMANDOLA L.J., TIELENS A.G.G.M. Astrophys. J. Suppl. Ser. 86, 713 (1993) Mid- and far-infrared spectroscopy of ices: optical constants and integrated absorbances

J/ApJS/87/367 See V/92A/

J/ApJS/87/451 See VII/158/

J/ApJS/88/423/  table1.dat table2.dat table3.dat  (40Kb)

Hui X., Ford H.C., Ciardullo R., Jacoby G.H. Astrophys. J. Suppl. Ser. 88, 423 (1993) The planetary nebula system and dynamics of NGC 5128. II. A planetary nebula catalog

J/ApJS/88/529 See VII/156/

J/ApJS/89/1 See VII/157/

J/ApJS/89/57/  table5.dat  (86Kb)

Yamada T., Takata T., Djamaluddin T., Tomita A., Aoki K., Takeda A., Saito M. Astrophys. J. Suppl. Ser. 89, 57 (1993) A search for IRAS galaxies behind the southern Milky Way.

J/ApJS/89/85/  table2 table3 table4 table5 table6 table7  (173Kb)

Ivanov G.R., Freedman W.L., Madore B.F. Astrophys. J. Suppl. Ser. 89, 85 (1993) A catalog of blue and red supergiants in M33.

J/ApJS/89/189/  table1 table2  (40Kb)

Chengalur J.N., Lewis B.M., Eder J., Terzian Y. Astrophys. J. Suppl. Ser. 89, 189 (1993) New OH/IR stars from color-selected IRAS sources. III. A complete survey

J/ApJS/89/259/  table1  (45Kb)

SCHMIDT E.G., CARRUTHERS G.R. Astrophys. J. Suppl. Ser. 89, 259 (1993) Far-ultraviolet stellar photometry: a field in Orion

J/ApJS/90/83/  table2 table3 table4  (198Kb)

Cote P., Welch D.L., Fischer P., Da Costa G.S., Tamblyn P., Seitzer P. Irwin M.J. Astrophys. J. Suppl. Ser. 90, 83 (1994) (From AAS CD-ROM series, Volume 2, 1994;) A search for binaries in the globular cluster NGC 3201

J/ApJS/90/173/  b1950.dat j2000.dat  (2.4 Mb)

Gregory P.C., Vavasour J.D., Scott W.K., Condon J.J. Astrophys. J. Suppl. Ser. 90, 173 (1994) (From AAS CD-ROM series, Volume 2, 1994) The Parkes-MIT-NRAO (PMN) map catalog of radio sources covering -88^{∘}dec. -37^{∘}at 4.85 GHz

J/ApJS/90/179 See VIII/27/

J/ApJS/91/111 See VIII/27/

J/ApJS/91/347/  table2.dat table4a.dat table4b.dat table4c.dat table4d.dat  (164Kb)

Becker R.H., White R.L., Helfand D.J., Zoonematkermani S. Astrophys. J. Suppl. Ser. 91, 347 (1994) A 5GHz VLA survey of the galactic plane.

J/ApJS/91/461/  notes3 spectrum table1 table2 table3 table4 table5 table6  (1.3 Mb)

THOMAS R.J., NEUPERT W.M. Astrophys. J. Suppl. Ser. 91, 461 (1994) (From AAS CD-ROM series, Volume 2, 1994) Extreme ultraviolet spectrum of a solar active region from SERTS

J/ApJS/91/507/  dina4p.sty table2 table2.tex  (238Kb)

DE JUAN L., COLINA L., PEREZ-FOURNON I. Astrophys. J. Suppl. Ser. 91, 507 (1994) (From AAS CD-ROM series, Volume 2, 1994) Surface photometry of low-luminosity radio galaxies

J/ApJS/91/583/  tables  (418Kb)

HILL R.J., MADORE B.F., FREEDMAN W.L. Astrophys. J. Suppl. Ser. 91, 583 (1994) (From AAS CD-ROM series, Volume 2, 1994) The initial mass function for massive stars in the Magellanic Clouds. I. UBV photometry and color-magnitude diagrams for 14 OB associations

J/ApJS/91/625/ table2.dat table2.tex table3.dat table3.tex table7.dat table7.tex table8.dat table8.tex (148Kb) Stauffer J.R., Caillault J.-P., Gagne M., Prosser C.F., Hartmann L.W. Astrophys. J. Suppl. 91, 625 (1994) A deep imaging survey of the Pleiades with ROSAT.

J/ApJS/91/659/  table3  (27Kb)

Kurtz S., Churchwell E., Wood D.O.S. Astrophys. J. Suppl. Ser. 91, 659 (1994)

Ultracompact HII regions II. New high-resolution radio images.

J/ApJS/92/53/ table1.dat table2.dat table3a.dat table3b.dat table3c.dat table5a.dat table5b.dat table6.dat (334Kb) Wilkes B., Tananbaum H., Worrall D.M., Avni Y., Oey M.S., Flanagan J. Astrophys. J. Suppl. Ser. 92, 53 (1994) (From AAS CD-ROM series, Volume 2, 1994) The Einstein database of IPC X-ray observations of optically selected and radio-selected quasars. I.

J/ApJS/92/125/  table.tex table3 table4 table5  (105Kb)

VASSILIADIS E., WOOD P.R. Astrophys. J. Suppl. Ser. 92, 125 (1994) (From AAS CD-ROM series, Volume 2, 1994) Post-asymptotic giant branch evolution of low- to intermediate-mass stars

J/ApJS/92/219/  table1 table2  (42Kb)

van den Bergh S. Astrophys. J. Suppl. Ser. 92, 219 (1994) A catalog of recent supernovae.

J/ApJS/93/145/  table2 table3  (257Kb)

KOLLGAARD R.I., BRINKMANN W., MCMATH CHESTER M., FEIGELSON E.D., HERTZ P., REICH P., WIELEBINSKI R. Astrophys. J. Suppl. Ser. 93, 145 (1994) (From AAS CD-ROM series, Volume 3, 1995) A Sensitive 1.5 GHz radio survey around the North Ecliptic Pole

J/ApJS/93/161/ table2 table3 table4 table5 table6 table7 table8 table9 table10 table11 table12 (87Kb) Sarajedini A., Norris J.E. Astrophys. J. Suppl. Ser. 93, 161 (1994) CCD photometry for six metal-rich galactic globular clusters.

J/ApJS/93/187/  table1 table1.tex  (91Kb)

Beauchamp A., Moffat A.F.J., Drissen L. Astrophys. J. Suppl. Ser. 93, 187 (1994) The galactic open cluster NGC 7419 and its five red supergiants.

J/ApJS/93/211/  table1 table2  (75Kb)

Diplas A., Savage B.D. Astrophys. J. Suppl. Ser. 93, 211 (1994) An IUE survey of interstellar H I Lyαabsorption. I. Column densities.

J/ApJS/93/411 See VIII/26/

J/ApJS/93/441 See VIII/26/

J/ApJS/93/549/  table1 table2 table3  (28Kb)

Lewis B.M. Astrophys. J. Suppl. Ser. 93, 549 (1994) Supplement to the Arecibo 1612 MHz survey of color-selected IRAS sources

J/ApJS/93/569/  euve.txt table1 table2 table7  (0Kb)

Bowyer S., Lieu R., Lampton M., Lewis J., Wu X., Drake J.J., Malina R.F. Astrophys. J. Suppl. Ser. 93, 569 (1994) The first Extreme Ultraviolet Explorer source catalog

J/ApJS/94/127/  refs.dat table1.dat  (120Kb)

Fruscione A., Hawkins I., Jelinsky P., Wiercigroch A. Astrophys. J. Suppl. Ser. 94, 127 (1994) The distribution of neutral hydrogen in the interstellar medium. I. The data.

J/ApJS/94/221/  table1 table2 table4 table5  (1.5 Mb)

Nave G., Johansson S., Learner R.C.M., Thorne A.P., Brault J.W. Astrophys. J. Suppl. Ser. 94, 221 (1994) (From AAS CD-ROM series, Volume 3, 1995) A new multiplet table for FeI

J/ApJS/94/461/  table3.dat table5.dat table6.dat  (40Kb)

Songaila A., Cowie L.L., Hu E.M., Gardner J.P. Astrophys. J. Suppl. Ser. 94, 461 (1994) The Hawai K-band survey. III. Spectroscopy of K 20 galaxies.

J/ApJS/94/615/ (0.0 Mb) HODAPP K.-W. Astrophys. J. Suppl. Ser. 94, 615 (1994) (From AAS CD-ROM series, Volume 3, 1995) A K’ imaging survey of molecular outflow sources

J/ApJS/94/687/  tablea1.dat tablea1.doc  (142Kb)

WORTHEY G., FABER S.M., GONZALEZ J.J., BURSTEIN D. Astrophys. J. Suppl. Ser. 94, 687 (1994) (From AAS CD-ROM series, Volume 3, 1995) Old stellar populations. V. Absorption feature indices for the complete Lick/IDS sample of stars

J/ApJS/94/749/  table1.dat table5.dat  (28Kb)

Kirkpatrick J.D., McGraw J.T., Hess T.R., Liebert J., McCarthy Jr D.W. Astrophys. J. Suppl. Ser. 94, 749 (1994) The luminosity function at the end of the main sequence: results of a deep, large-area, CCD survey for cool dwarfs.

J/ApJS/95/1/ table3.tex table3a table3b table6 table6.tex table9 table9.tex table10 table10.tex table11 table11.tex table12 table12.tex tablea1 tablea2 tablea3 tablea4 tablea5 tableb1 (865Kb) Elvis M., Wilkes B.J., McDowell J.C., Green R.F., Bechtold J., Willner S.P., Oey M.S., Polomski E., Cutri R. Astrophys. J. Suppl. Ser. 95, 1 (1994) (From AAS CD-ROM series, Volume 3, 1995 Lee Brotzman [ADS] 13-Nov-94) Atlas of Quasar Energy Distributions

J/ApJS/95/163/ hr1175.ps hr1206.ps hr1240.ps hr1335.ps hr1338.ps hr1371.ps hr1400.ps hr1530.ps hr1550.ps hr1640.ps hr1670.ps hr1718.ps hr1858.ps hr2586.ps hr2800.ps hrall.ps sp/* table2.dat table2.tex table3.dat table3.tex table4.dat table4.tex (5.0 Mb) Snow T.P., Lamers H.J.G.L.M., Lindholm D.M., Odell A.P. Astrophys. J. Suppl. Ser. 95, 163 (1994) An atlas of ultraviolet P Cygni profiles

J/ApJS/95/345/  fig/* table1.dat table2 table3.dat table4  (17.9 Mb)

Taylor G.B., Vermeulen R.C., Pearson T.J., Readhead A.C.S., Henstock D.R., Browne I.W.A., Wilkinson P.N. Astrophys. J. Suppl. Ser. 95, 345 (1994) (From AAS CD-ROM series, Volume 4, 1995 Lee Brotzman [ADS] 07-May-95) The Second Caltech-Jodrell Bank VLBI Survey. I. Observations of 91 of 193 Sources

J/ApJS/95/387/  note2 table2  (25Kb)

Zamorano J., Rego M., Gallego J., Vitores A.G., Gonzalez-Riestra R., Rodriguez-Caderot G. Astrophys. J. Suppl. Ser. 95, 387 (1994) Survey of emission-line galaxies: Universidad Complutense de Madrid list 1.

J/ApJS/95/401/  table1a.dat table1b.dat table1c.dat table1d.dat  (24Kb)

Plionis M. Astrophys. J. Suppl. Ser. 95, 401 (1994) Position angles and alignments of clusters of galaxies

J/ApJS/95/419/  table1 table2  (56Kb)

Dobashi K., Bernard J.-P., Yonekura Y., Fukui Y. Astrophys. J. Suppl. Ser. 95, 419 (1994) Molecular clouds in Cygnus. I. A large-scale ^{13}CO survey.

J/ApJS/95/457/  table2 table3 table4 table5  (51Kb)

Wood D.O.S., Myers P.C., Daugherty D.A. Astrophys. J. Suppl. Ser. 95, 457 (1994) IRAS images of nearby dark clouds

J/ApJS/95/535/ readme.doc table6.dat table6.tex table7.dat table7.tex table8.dat table8.tex table9.dat table9.tex table10.dat table10.tex table11.dat table11.tex table12.dat table13.dat table14.dat table15.dat (10.1 Mb) Goorvitch D. Astrophys. J. Suppl. Ser. 95, 535 (1994) (From AAS CD-ROM series, Volume 3, 1995) Infrared CO line list for the X ^{1}Σ^{+} state

J/ApJS/96/39/  csrg.dat notes.dat table6.dat  (512Kb)

Buta R. Astrophys. J. Suppl. Ser. 96, 39 (1995) The catalog of southern ringed galaxies.

J/ApJS/96/117/  table3.dat table4.dat table5.dat  (28Kb)

McLeod B.A., Bernstein G.M., Rieke M.J., Tollestrup E.V., Fazio G.G. Astrophys. J. Suppl. Ser. 96, 117 (1995) K-Band Galaxy Counts

J/ApJS/96/123/  table2 table4  (150Kb)

Tormen G., Burstein D. Astrophys. J. Suppl. Ser. 96, 123 (1995)

Recalibration of the H_{−0.5}-magnitudes of spiral galaxies

J/ApJS/96/175/  table2.dat  (139Kb)

Beers T.C., Sommer-Larsen J. Astrophys. J. Suppl. Ser. 96, 175 (1995) (From AAS CD-ROM series, Volume 4, 1995 Lee Brotzman [ADS] 08-May-95) Kinematics of Metal-Poor Stars in the Galaxy

J/ApJS/96/303/  table1  (20Kb)

Weaver K.A., Arnaud K.A., Boldt E.A., Christian D., Corcoran M., Holt S.S., Jahoda K., Kelley R., Marshall F.E., Mushotzky R.F., Petre R., Rawley G., Serlemitsos P.J., Schlegel E.M., Smale A.P., Swank J.H., Szymkowiak A.E. Astrophys. J. Suppl. Ser. 96, 303 (1995) Calibrating the broad band X-ray telescope.

J/ApJS/96/343/  table2 table3  (56Kb)

Quintana H., Ramirez A. Astrophys. J. Suppl. Ser. 96, 343 (1995) Redshifts of 165 Abell and southern rich clusters of galaxies.

J/ApJS/96/461/  table2.dat  (307Kb)

Bowyer S., Sasseen T.P., Wu X., Lampton M. Astrophys. J. Suppl. Ser. 96, 461 (1995) (From AAS CD-ROM series, Volume 4, 1995 Lee Brotzman [ADS] 08-May-95) A Catalog of Far-Ultraviolet Point Sources Detected with the FAUST Telescope on ATLAS-1

J/ApJS/96/605/  table4 table5  (87Kb)

Schmidt E.G., Carruthers G.R. Astrophys. J. Suppl. Ser. 96, 605 (1995) (From AAS CD-ROM series, Volume 4, 1995 Lee Brotzman [ADS] 08-May-95) Far-Ultraviolet Stellar Photometry: Fields in Sagittarius and Scorpius

J/ApJS/97/1/  notes.dat table1.dat  (22Kb)

Lockman F.J., Savage B.D. Astrophys. J. Suppl. Ser. 97, 1 (1995) The Hubble Space Telescope Quasar Absorption Line Key Project. X. Galactic H I 21 centimeter emission toward 143 quasars and active galactic nuclei.

J/ApJS/97/77/  table1  (55Kb)

Doi M., Fukugita M., Okamura S., Tarusawa K. Astrophys. J. Suppl. Ser. 97, 77 (1995) (From AAS CD-ROM series, Volume 4, 1995 Lee Brotzman [ADS] 08-May-95) Automated Surface Photometry for the Coma Cluster Galaxies: The Catalog

J/ApJS/97/141/  table1a table1b tables.tex  (178Kb)

Eskridge P.B., Fabbiano G., Kim D.-W. Astrophys. J. Suppl. Ser. 97, 141 (1995) A multiparametric analysis of the Einstein sample of early-type galaxies. I. Luminosity and ISM parameters.

J/ApJS/97/189 See V/92A/

J/ApJS/97/347 See VIII/38/

J/ApJS/98/1/  fig/* table1.dat table2.dat table4.dat table5  (29.7 Mb)

Polatidis A.G., Wilkinson P.N., Xu W., Readhead A.C.S., Pearson T.J., Taylor G.B., Vermeulen R.C. Astrophys. J. Suppl. Ser. 98, 1 (1995) (From AAS CD-ROM series, Volume 4, 1995 Lee Brotzman [ADS] 10-May-95) The First Caltech-Jodrell Bank VLBI survey. I. Lambda = 18 centimeter observations of 87 sources

J/ApJS/98/33/  fig/* table2.dat table3  (9.8 Mb)

Thakkar D.D., Xu W., Readhead A.C.S., Pearson T.J., Taylor G.B., Vermeulen R.C., Polatidis A.G., Wilkinson P.N. Astrophys. J. Suppl. Ser. 98, 33 (1995) (From AAS CD-ROM series, Volume 4, 1995 Lee Brotzman [ADS] 10-May-95) The First Caltech-Jodrell Bank VLBI Survey. II. Lambda = 18 Centimeter Observations of 25 Sources

J/ApJS/98/369/  appen table2a table2b table3 table4  (136Kb)

Bicay M.D., Kojoian G., Seal J., Dickinson D.F., Malkan M.A. Astrophys. J. Suppl. Ser. 98, 369 (1995) A multifrequency radio continuum and IRAS faint source survey of Markarian galaxies

J/ApJS/98/477/  table2  (66Kb)

Ho L.C., Filippenko A.V., Sargent W.L. Astrophys. J. Suppl. Ser. 98, 477 (1995) A search for ”dwarf” Seyfert nuclei. II. An optical spectral atlas of the nuclei of nearby galaxies

J/ApJS/98/595/ aastab.sty environ.tex fig/* table2.dat table2.tex table3.dat table3.tex table4.dat table4.tex (95.9 Mb) Hill J.K., Isensee J.E., Bohlin R.C., Cheng K.P., Hintzen P.M.N., OConnell R.W., Roberts M.S., Smith A.M.,Smith E.P.,Stecher T.P. Astrophys. J. Suppl. Ser. 98, 595 (1995) (From AAS CD-ROM series, Volume 5, 1995 Lee Brotzman [ADS] 02-Aug-95) Ultraviolet and optical observations of OB associations in M31

J/ApJS/98/659/  table1 table1.tex table3 table3.tex  (144Kb)

Zhang C.Y. Astrophys. J. Suppl. Ser. 98, 659 (1995) A statistical distance scale for galactic planetary nebulae

J/ApJS/99/135/  table2  (120Kb)

Abt H.A., Morrell N.I. Astrophys. J. Suppl. Ser. 99, 135 (1995) (From AAS CD-ROM series, Volume 5, 1995 Lee Brotzman [ADS] 02-Aug-95) The relation between rotational velocities and spectral peculiarities among A-type stars.

J/ApJS/99/173/  models.all  (5.4 Mb)

Leitherer C., Robert C., Heckman T.M. Astrophys. J. Suppl. Ser. 99, 173 (1995) (From AAS CD-ROM series, Volume 5, 1995 Lee Brotzman [ADS] 29-Aug-95) Atlas of synthetic ultraviolet spectra of massive star populations

J/ApJS/99/297/  table2 table3 table4.dat table5.dat  (37Kb)

Xu W., Readhead A.C.S., Pearson T.J., Polatidis A.G., Wilkinson P.N. Astrophys. J. Suppl. Ser. 99, 297 (1995) (From AAS CD-ROM series, Volume 5, 1995 Lee Brotzman [ADS] 29-Aug-95) The First Caltech-Jodrell Bank VLBI Survey. III. VLBI and Merlin Observations at 5 GHz and VLA Observations at 1.4 GHz

J/ApJS/99/391/  table1-2  (0Kb)

Huchra J.P., Geller M.J., Corwin H.G.Jr Astrophys. J. Suppl. Ser. 99, 391 (1995) The CfA redshift survey: data for the NGP +36 zone.

J/ApJS/99/501/  table1.dat table2.dat  (92Kb)

Persic M., Salucci P. Astrophys. J. Suppl. Ser. 99, 501 (1995) Rotation curves of 967 spiral galaxies

J/ApJS/99/551/  table2.dat table3.dat  (165Kb)

Hunter D.A., Boyd D.M., Hawley W.N. Astrophys. J. Suppl. Ser. 99, 551 (1995) (From AAS CD-ROM series, Volume 5, 1995 Lee Brotzman [ADS] 31-Aug-95) The massive star populations in ionized gas shells in M33 and M31

J/ApJS/99/637/  table.tex table1 table1.tex table2 table2.tex table3 table3.tex  (308Kb)

Norris J.E., Da Costa G.S., Tingay S.J. Astrophys. J. Suppl. Ser. 99, 637 (1995) (From AAS CD-ROM series, Volume 5, 1995 Lee Brotzman [ADS] 31-Aug-95) Equivalent widths for 54 red giants in 47 Tucanae, Omega Centauri, NGC 6397, M22, and NGC 6752

J/ApJS/99/701/  table1  (21Kb)

Fleming T.A., Molendi S., Maccacaro T., Wolter A. Astrophys. J. Suppl. Ser. 99, 701 (1995) THE EINSTEIN Extended medium sensitivity survey second epoch: results for the stars

J/ApJS/100/1/  fig/* table2.dat table3  (16.7 Mb)

Henstock D.R., Browne I.W.A., Wilkinson P.N., Taylor G.B., Vermeulen R.C., Pearson T.J., Readhead A.C.S. Astrophys. J. Suppl. Ser. 100, 1 (1995) (From AAS CD-ROM series, Volume 4, 1995 Lee Brotzman [ADS] 10-May-95) The Second Caltech-Jodrell Bank VLBI Survey. II. Observations of 102 of 193 Sources

J/ApJS/100/69 See VII/185/

J/ApJS/100/105/  table1.dat  (39Kb)

McElroy D.B. Astrophys. J. Suppl. Ser. 100, 105 (1995) A catalog of stellar velocity dispersions. II. 1994 update

J/ApJS/100/125/ b-0d25.dat b-0d50.dat b-0d75.dat b-1d00.dat b0d00.dat b0d25.dat b0d50.dat b0d75.dat b1d00.dat (4.4 Mb) Sakamoto S., Hasegawa T., Hayashi M., Handa T., Oka T. Astrophys. J. Suppl. Ser. 100, 125 (1995) An Out-of-Plane CO (J = 2-1) Survey of the Milky Way. I. The Data

J/ApJS/100/281/ table1.dat table2.dat table3.dat table4.dat table5.dat table6.dat table7.dat table8.dat table9.dat table10.dat table11.dat table12.dat table13.dat (2.8 Mb) Sugiyama N. Astrophys. J. Suppl. Ser. 100, 281 (1995) (From AAS CD-ROM series, Volume 5, 1995 Lee Brotzman [ADS] 25-Sep-95) Cosmic background anisotropies in cold dark matter cosmology

J/ApJS/101/41/  table2.dat table3.dat  (147Kb)

Bica E.L.D., Schmitt H.R. Astrophys. J. Suppl. Ser. 101, 41 (1995) A revised and extended catalog of Magellanic system clusters, associations, and emission nebulae. I. Small Magellanic Cloud and Bridge.

J/ApJS/101/117/  table.tex tablea1.dat tablea2.dat tablea3.dat tablea4.dat tablea5.dat  (240Kb)

Kenyon S.J., Hartmann L. Astrophys. J. Suppl. Ser. 101, 117 (1995) (From AAS CD-ROM series, Volume 5, 1995 Lee Brotzman [ADS] 06-Nov-95) Pre-main-sequence evolution in the Taurus-Auriga molecular cloud

J/ApJS/101/259/  table4.dat  (0Kb)

Thompson D.J., Bertsch D.L., Dingus B.L., Esposito J.A., Etienne A., Fichtel C.E., Friedlander D.P., Hartman R.C., Hunter S.D., Kendig D.J., Mattox J.R., McDonald L.M., von Montigny C., Mukherjee R., Ramanamurthy P.V., Sreekumar P., Fierro J.M., Lin Y.C., Michelson P.F., Nolan P.L., Shriver S.K., Willis T.D., Kanbach G., Mayer-Hasselwander H.A., Merck M., Radecke H.-D., Kniffen D.A., Schneid E.J. Astrophys. J. Suppl. Ser. 101, 259 (1995) The second EGRET catalog of high-energy gamma-ray sources

J/ApJS/101/287/  galaxies hii.dat  (153Kb)

Tsvetanov Z.I., Petrosian A.R. Astrophys. J. Suppl. Ser. 101, 287 (1995) (From AAS CD-ROM series, Volume 6, 1996 Lee Brotzman [ADS] 23-Apr-96) H II region populations in Seyfert galaxies. I. Database

J/ApJS/102/29/  table1 table2 table3 table3b table4  (122Kb)

Huchra J.P., Brodie J.P., Caldwell N., Christian C., Schommer R. Astrophys. J. Suppl. Ser. 102, 29 (1996) Extragalactic globular clusters. IV. The data.

J/ApJS/102/75/  refs.dat table1.dat table5.dat  (38Kb)

Micela G., Sciortino S., Kashyap V., Harnden F.R., Rosner R. Astrophys. J. Suppl. 102, 75 (1996) ROSAT Observations of the Pleiades. I. X-Ray Characteristics of a Coeval Stellar Population

J/ApJS/102/129 See II/203A/

J/ApJS/102/289/  table2 table2.tex  (88Kb)

Yee H.K.C., Ellington E., Abraham R.G., Gravel P., Carlberg R.G., Smecker-Hane T.A., Schade D., Rigler M. Astrophys. J. Suppl. Ser. 102, 289 (1996) The CNOC cluster redshift survey catalogs. II. Abell 2390.

J/ApJS/103/145 See VIII/38/

J/ApJS/103/235/ ab10750.dat ab10850.dat ab10950.dat ab11050.dat ab11150.dat ab11250.dat ab11350.dat ab11450.dat linelist.dat (567Kb) Wallace L., Hinkle K. Astrophys. J. Suppl. Ser. 103, 235 (1996) (From AAS CD-ROM series, Volume 6, 1996 Lee Brotzman [ADS] 23-Apr-96) An infrared spectral atlas of Arcturus for the range 10750-11500cm-1 (8690-9300A)

J/ApJS/103/331/  table1.dat table2.dat table3.dat  (37Kb)

Lu N.Y., Hoffman G.L., Salpeter E.E., Houck J.R. Astrophys. J. Suppl. Ser. 103, 331 (1996) Radio properties and optical identification of two samples of milliJansky radio sources at 1.4 GHz

J/ApJS/103/427 See VIII/40/

J/ApJS/103/433/  table3.dat  (216Kb)

Beers T.C., Wilhelm R., Doinidis S.P., Mattson C.J. Astrophys. J. Suppl. Ser. 103, 433 (1996) (From AAS CD-ROM series, Volume 6, 1996 Lee Brotzman [ADS] 03-May-96) A catalog of candidate field horizontal-branch and A-type stars. II.

J/ApJS/104/71/ figure1a.ps figure1b.ps figure1c.ps figure1d.ps figure1e.ps figure1f.ps figure1g.ps figure1h.ps figure1i.ps figure1j.ps figure1k.ps figure1l.ps figure1m.ps table4a.dat table4b.dat table4c.dat table4d.dat table4e.dat table4f.dat table4g.dat table4h.dat table4i.dat table4j.dat table4k.dat table4l.dat table4m.dat (32.9 Mb) Oey M.S. Astrophys. J. Suppl. Ser. 104, 71 (1996) (From AAS CD-ROM series, Volume 6, 1996 Lee Brotzman [ADS] 03-May-96) UBV photometry of OB associations within superbubbles of the Large Magellanic Cloud

J/ApJS/104/101/  table2 table3 table5 table5.fit  (308Kb)

Schmidt E.G., Carruthers G.R. Astrophys. J. Suppl. Ser. 104, 101 (1996) Far-ultraviolet stellar photometry: fields centered on Rho Ophiuchi and the Galactic center

J/ApJS/104/185/  table1 table2 tables.tex  (28Kb)

Hall P.B., Osmer P.S., Green R.F., Porter A.C., Warren S.J. Astrophys. J. Suppl. Ser. 104, 185 (1996) A deep multicolor survey. I. Imaging observations and catalog of stellar objects.

J/ApJS/104/199 See VII/200/

J/ApJS/104/287/  table1.dat table1.tex table2.dat table2.tex  (393Kb)

Smith E.P., Pica A.J., Bohlin R.C., Cornett R.H., Fanelli M.N., Landsman W.B., O’Connell R.W., Roberts M.S., Smith A.M., Stecher T.P. Astrophys. J. Suppl. Ser. 104, 287 (1996) (From AAS CD-ROM Vol.7 Lee Brotzman [ADS] 16-Jul-96) Ultraviolet imaging telescope near-ultraviolet bright object catalog.

J/ApJS/105/93/  table2.dat  (191Kb)

Evans I.N., Koratkar A.P., Storchi-Bergmann T., Kirkpatrick H., Heckman T.M., Wilson A.S. Astrophys. J. Suppl. Ser. 105, 93 (1996) (From AAS CD-ROM series, Volume 7, 1996 Lee Brotzman [ADS] 16-Jul-96) An Atlas of H II Regions in Nearby Seyfert Galaxies

J/ApJS/105/209/  table1.dat table2.dat table3.dat  (78Kb)

Impey C.D., Sprayberry D., Irwin M.J., Bothun G.D. Astrophys. J. Suppl. Ser. 105, 209 (1996) Low surface brightness galaxies in the local universe. I. The catalog.

J/ApJS/105/343/  note2 table2 table3 tables.tex  (61Kb)

Zamorano J., Gallego J., Rego M., Vitores A.G., Alonso O. Astrophys. J. Suppl. Ser. 105, 343 (1996) Survey of emission-line galaxies: Universidad Complutense de Madrid list 2.

J/ApJS/105/369/  data.fit table1.dat  (1.3 Mb)

Murphy E.M., Lockman F.J., Laor A., Elvis M. Astrophys. J. Suppl. Ser. 105, 369 (1996) (From AAS CD-ROM series, Volume 7, 1996 Lee Brotzman [ADS] 22-Oct-96) Galactic H I column densities toward quasars and active galactic nuclei

J/ApJS/105/419/  sp/* table4.dat  (1.2 Mb)

Barnbaum C., Stone R.P.S., Keenan P.C. Astrophys. J. Suppl. 105, 419 (1996) A Moderate-Resolution Spectral Atlas of Carbon Stars: R, J, N, CH, and Barium Stars

J/ApJS/105/475/  table1.dat table2.dat  (48Kb)

Heintz W.D. Astrophys. J. Suppl. Ser. 105, 475 (1996) Observations of double stars and new pairs. XVII.

J/ApJS/106/1/  cluster.dat table3.dat table4.dat  (66Kb)

Wegner G., Colless M., Baggley G., Davies R.L., Bertschinger E., Burstein D., McMahan R.K.Jr, Saglia R.P. Astrophys. J. Suppl. Ser. 106, 1 (1996) The peculiar motions of early-type galaxies in two distant regions. I. Cluster and galaxy selection.

J/ApJS/106/27/  table2a table2b  (155Kb)

Keel W.C. Astrophys. J. Suppl. Ser. 106, 27 (1996) (From AAS CD-ROM series, Volume 7, 1996 Lee Brotzman [ADS] 09-Oct-96) Rotation curves and velocity measures for spiral galaxies in pairs

J/ApJS/106/399/  refs table1 table2 table3 table4 table5  (109Kb)

Polletta M., Bassani L., MalagutI G., Palumbo G.G.C., Caroli E. Astrophys. J. Suppl. Ser. 106, 399 (1996) A multiwavelength catalog of Seyfert 2 galaxies observed in the 2-10 keV energy band

J/ApJS/106/463/  table1a table2a table3 tables.tex  (62Kb)

Jiang B.W., Deguchi S., Yamamura I., Nakada Y., Cho S.H., Yamagata T. Astrophys. J. Suppl. Ser. 106, 463 (1996) (From AAS CD-ROM series, Volume 7, 1996 Lee Brotzman [ADS] 09-Nov-96) SiO maser sources in the outer disk of the Galaxy

J/ApJS/107/1/  notes.dat table1.dat  (52Kb)

Abraham R.G., van den Bergh S., Glazebrook K., Ellis R.S., Santiago B.X., Surma P., Griffiths R.E. Astrophys. J. Suppl. Ser. 107, 1 (1996) The morphologies of distant galaxies. II. Classifications from the Hubble space telescope medium deep survey.

J/ApJS/107/19/  notes.dat table2.dat  (58Kb)

McCarthy P.J., Kapahi V.K., Van Breugel W., Persson S.E., Athreya R., Subrahmanya C.R. Astrophys. J. Suppl. Ser. 107, 19 (1996) The Molonglo Reference Catalog/1 Jansky radio source survey. I. Radio galaxy identifications.

J/ApJS/107/97/  table1.dat  (537Kb)

Mathewson D.S., Ford V.L. Astrophys. J. Suppl. Ser. 107, 97 (1996) (From AAS CD-ROM series, Volume 7, 1996 Lee Brotzman [ADS] 04-Nov-96) Parameters of 2447 southern spiral galaxies for use in the Tully-Fisher relation

J/ApJS/107/201/  fields.dat galaxy.dat holes.dat  (263Kb)

Loveday J., Peterson B.A., Maddox S.J., Efstathiou G. Astrophys. J. Suppl. Ser. 107, 201 (1996) The Stromlo-APM Redshift Survey IV: The Redshift Catalogue

J/ApJS/107/215/  atlas.dat  (29Kb)

Maoz D., Filippenko A. V., Ho L. C., Macchetto D., Rix H., Schneider D. P. Astrophys. J. Suppl. 107, 215 (1996) An Atlas of Hubble Space Telescope Ultraviolet Images of Nearby Galaxies

J/ApJS/107/227/ catsupp.txt figure1.ps figure2.ps figure3.ps figure4.ps figure5.ps figure6.ps figure7.ps figure8.ps figure9.ps figure10.ps table1.dat table2 (1.1 Mb) Thompson D.J., Bertsch D.L., Dingus B.L., Esposito J.A., Etienne A., Fichtel C.E., Friedlander D.P., Hartman R.C., Hunter S.D., Kendig D.J., Mattox J.R., McDonald L.M., Mukherjee R., Ramanamurthy P.V., Sreekumar P., Von Montigny C., Fierro J.M., Jones B.B., Lin Y.C., Michelson P.F., Nolan P.L., Tompkins W., Willis T.D., Kanbach G., Mayer-Hasselwander H.A., Merck M., Pohl M., Kniffen D.A., , Schneid E.J. Astrophys. J. Suppl. Ser. 107, 227 (1996) (From AAS CD-ROM series, Volume 8, 1997 Lee Brotzman [ADS] 27-Feb-97) Supplement to the Second EGRET Catalog of High-Energy Gamma-Ray Sources

J/ApJS/107/239/  table1.dat wsrt.dat  (348Kb)

Taylor A.R., Goss W.M., Coleman P.H., van Leeuwen J., Wallace B.J. Astrophys. J. Suppl. Ser. 107, 239 (1996) A Westerbork synthesis radio telescope 327 MHz survey of the galactic plane.

J/ApJS/107/281/  fits/* table1.dat table1.tex  (40Kb)

Hanson M.M., Conti P.S., Rieke M.J. Astrophys. J. Suppl. Ser. 107, 281 (1996) (From AAS CD-ROM series, Volume 7, 1996 Lee Brotzman [ADS] 09-Oct-96) A spectral atlas of hot, luminous stars at 2 microns

J/ApJS/107/521/  table1.dat  (48Kb)

Visvanathan N., Yamada T.

Astrophys. J. Suppl. Ser. 107, 521 (1996) (From AAS CD-ROM series, Volume 8, 1997 Lee Brotzman [ADS] 06-May-97) Redshift distribution of galaxies in the southern Milky way region 210^{∘}l 360^{∘}and —b— 15^{∘}.

J/ApJS/108/99 See VII/205/

J/ApJS/108/545 See II/203A/

J/ApJS/109/45/  table5.dat table6.dat  (128Kb)

Munn J.A., Koo D.C., Kron R.G., Majewski S.R., Bershady M.A., Smetanka J.J. Astrophys. J. Suppl. Ser. 109, 45 (1997) The Kitt Peak galaxy redshift survey with multicolor photometry: basic data

J/ApJS/109/333 See VII/198/

J/ApJS/109/367/  table4  (18Kb)

Knopp G.P., Chambers K.C. Astrophys. J. Suppl. Ser. 109, 367 (1997) Deep imaging of the field of the high-redshift radio source 4C 23.56

J/ApJS/109/489/  table1.dat table2.dat  (36Kb)

Lewis B.M. Astrophys. J. Suppl. Ser. 109, 489 (1997) Main-line OH observations of the Arecibo set of OH/IR stars

J/ApJS/110/1/  table2  (25Kb)

de Carvalho R.R., Ribeiro A.L.B., Capelato H.V., Zepf S.E. Astrophys. J. Suppl. 110, 1 (1997) Redshift survey of galaxies around a selected sample of compact groups.

J/ApJS/110/21/  refs.dat table1.dat table2.dat  (33Kb)

Yonekura Y., Dobashi K., Mizuno A., Ogawa H., Fukui Y. Astrophys. J. Suppl. Ser. 110, 21 (1997) Molecular clouds in Cepheus and Cassiopeia.

J/ApJS/110/213/ imagebc.fit imageia.fit imageic.fit imageie.fit imageih.fit imageij.fit imageik.fit imagerb.fit imagerd.fit imagerf.fit imagerg.fit imageri.fit imageva.fit imageve.fit imagevh.fit imagevj.fit imagevk.fit plate1a.ps plate1b.ps plate1c.ps plate1d.ps plate1e.ps plate1f.ps plate1g.ps plate1h.ps plate1i.ps plate1j.ps plate1k.ps table1.dat table3.dat table4.dat (211.1 Mb) Smail I., Dressler A., Couch W.J., Ellis R.S., Oemler A.Jr., Butcher H., Sharples R.M. Astrophys. J. Suppl. Ser. 110, 213 (1997) A Catalog of Morphological Types in 10 Distant Rich Clusters of Galaxies

J/ApJS/110/227 See VII/205/

J/ApJS/111/143/  table1 table2 table4 table5  (45Kb)

Brosch N., Formiggini L., Almoznino E., Sasseen T., Lampton M., Bowyer S. Astrophys. J. Suppl. Ser. 111, 143 (1997) FAUST observations of ultraviolet sources toward the Virgo cluster

J/ApJS/111/163/ table5.dat table6.dat table7.dat table9.dat table10.dat table11.dat table12.dat table13.dat table14.dat (168Kb) Burstein D., Jones C., Forman W., Marston A.P., Marzke R.O. Astrophys. J. Suppl. Ser. 111, 163 (1997) (From AAS CD-ROM series, Volume 9, 1997, Lee Brotzman [ADS] 17-Jun-97) An Einstein X-ray survey of optically selected galaxies. I. Data

J/ApJS/111/203/  ngc3379.dat ngc4472.dat ngc4594.dat  (41Kb)

Vazdekis A., Peletier R., Beckman J.E., Casuso E. Astrophys. J. Suppl. Ser. 111, 203 (1997) (From AAS CD-ROM series, Volume 8, 1997 Lee Brotzman [ADS] 29-Apr-97) A New Chemo-evolutionary Population Synthesis Model for Early-Type Galaxies. II. Observations and Results

J/ApJS/111/377/  table2.dat  (45Kb)

Worthey G., Ottaviani D.L. Astrophys. J. Suppl. Ser. 111, 377 (1997) (From AAS CD-ROM series, Volume 9, 1997 Lee Brotzman [ADS] 25-Jul-97) Hγand Hδabsorption features in stars and stellar populations

J/ApJS/111/445 See III/240/

J/ApJS/112/29/  tables.dat  (54Kb)

Feinstein C.

Astrophys. J. Suppl. Ser. 112, 29 (1997) H II regions in southern spiral galaxies: the Hαluminosity function

J/ApJS/112/109/  tab3note.txt table1.dat table2.dat table3.dat table4.dat table5.dat  (496Kb)

Barsony M., Kenyon S.J., Lada E.A., Teuben P.J. Astrophys. J. Suppl. Ser. 112, 109 (1997) A Near-Infrared Imaging Survey of the Rho Ophiuchi Cloud Core

J/ApJS/112/245/  table5.dat table6.dat  (91Kb)

Nakanishi K., Takata T., Yamada T., Takeuchi T.T., Shiroya R., Miyazawa M., Watanabe S., Saito M. Astrophys. J. Suppl. Ser. 112, 245 (1997) Search and redshift survey for IRAS galaxies behind the Milky Way and structure of the Local Void

J/ApJS/112/315/ table1.dat table2.dat table3.dat table4.dat table7.dat table9.dat table10.dat table11.dat table12.dat table13.dat (279Kb) Ho L.C., Filippenko A.V., Sargent W.L.W. Astrophys. J. Suppl. Ser. 112, 315 (1997) A search for ”Dwarf” Seyfert nuclei. III. Spectroscopic parameters and properties of the host galaxies.

J/ApJS/112/487/  table2.dat table3.dat table4.dat table5.dat tables.tex  (55Kb)

Soker N. Astrophys. J. Suppl. Ser. 112, 487 (1997) Properties that cannot be explained by the progenitors of planetary nebulae

J/ApJS/112/557/  table1.dat table1.tsv  (2.0 Mb)

Kwok S., Polk K., Bidelman W.P. Astrophys. J. Suppl. Ser. 112, 557 (1997) (From AAS CD-ROM series, Volume 9, 1997 Lee Brotzman [ADS] 27-Jul-97) Classification and Identification of IRAS Sources With Low-Resolution Spectra

J/ApJS/113/1/  table2.dat table3.dat  (52Kb)

Ellingson E., Yee H.K.C., Abraham R.G., Morris S.L., Carlberg R.G., Smecker-Hane T.A. Astrophys. J. Suppl. Ser. 113, 1 (1997) The CNOC cluster redshift survey catalogs. III. MS 1621.5+2640 and MS 0302.7+1658

J/ApJS/113/121/  header.dat tables.dat  (89Kb)

Kato M. Astrophys. J. Suppl. Ser. 113, 121 (1997) (From AAS CD-ROM series, Volume 9, 1997 Lee Brotzman [ADS] 27-Aug-97) Optically thick winds from degenerate dwarfs. I. Classical nova of populations I and II

J/ApJS/115/1/  table1.dat table2.dat  (63Kb)

Slinglend K., Batuski D., Miller C., Haase S., Michaud K., Hill J.M. Astrophys. J. Suppl. Ser. 115, 1 (1998) The MX northern Abell cluster redshift survey.

J/ApJS/115/59/  table1.dat table2.dat table3.dat table5.dat  (56Kb)

Kun M. Astrophys. J. Suppl. Ser. 115, 59 (1998) Star formation in the Cepheus flare molecular clouds. I. Distance determination and the young stellar object candidates.

J/ApJS/115/185/ bflux.dat c5177.dat hablue.dat hacent.dat hared.dat hatot.dat hbblue.dat hbcent.dat hbred.dat hbtot.dat he4686.dat he5876.dat hgamma.dat iflux.dat rflux.dat vflux.dat (58Kb) Dietrich M., Peterson B.M., Albrecht P., Altmann M., Barth A.J., Bennie P.J., Bertram R., Bochkarev N.G., Bock H., Braun J.M., Burenkov A., Collier S., Fang L.-Z., Francis O.P., Filippenko A.V., Foltz C.B., Gaessler W., Gaskell C.M., Geffert M., Ghosh K.K., Hilditch R.W., Honeycutt R.K., Horne K., Huchra J.P., Kaspi S., Kuemmel M., Leighly K.M., Leonard D.C., Malkov Y.F., Mikhailov V., Miller H.R., Morrill A.C., Noble J., O’Brien P.T., Oswalt T.D., Pebley S.P., Pfeiffer M., Pronik V.I., Qian B.-C., Robertson J.W., Robinson A., Rumstay K.S., Schmoll J., Sergeev S.G., Sergeeva E.A., Shapovalova A.I., Skillman D.R., Snedden S.A., Soundararajaperumal S., Stirpe G.M., Tao J., Turner G.W., Wagner R.M., Wagner S.J., Wei J.Y., Wu H., Zheng W., Zou Z.L. Astrophys. J. Suppl. Ser. 115, 185 (1998) Steps toward determination of the size and structure of the broad-line region in active galactic nuclei. XII. Ground-based monitoring of 3C 390.3.

J/ApJS/115/271/  cross.dat intro.txt mk.dat notes.dat refs.dat ubvbeta.dat  (12Kb)

Reed B. Cameron Astrophys. J. Suppl. Ser. 115, 271 (1998) UBV-βDatabase for Case-Hamburg Northern and Southern Luminous Stars J/ApJS/116/1/ gal/* galcat.dat star/* starcat.dat table4.dat table7.dat table8.dat table9.dat table10.dat

(16.9 Mb) Trager S.C., Worthey G., Faber S.M., Burstein D., Gonzalez J.J. Astrophys. J. Suppl. Ser. 116, 1 (1998) Old Stellar Populations. VI. Absorption-Line Spectra of Galaxy Nuclei and Globular Clusters

J/ApJS/116/211/  table2.dat table3.dat  (44Kb)

Yee H.K.C., Ellingson E., Morris S.L., Abraham R.G., Carlberg R.G. Astrophys. J. Suppl. Ser. 116, 211 (1998) The CNOC cluster redshift survey catalogs. IV. MS 1358.4+6245 and MS 1008.1-1224

J/ApJS/116/231/  table2.dat table3.dat  (38Kb)

Abraham R.G., Yee H.K.C., Ellingson E., Carlberg R.G., Gravel P. Astrophys. J. Suppl. Ser. 116, 231 (1998) The CNOC cluster redshift survey catalogs. V. MS 1224.7+2007 and MS 1512.4+3647.

J/ApJS/116/247/  table2.dat table3.dat table4.dat  (42Kb)

Ellingson E., Yee H.K.C., Abraham R.G., Morris S.L., Carlberg R.G. Astrophys. J. Suppl. Ser. 116, 247 (1998) The CNOC cluster redshift survey catalogs. VI. MS 0015.9+1609 and MS 0451.5-0305

J/ApJS/116/263/ table4.dat table5.dat table6.dat table7.dat table8.dat table9.dat table10.dat table11.dat table12.dat (104Kb) Piatti A.E., Claria J.J., Bica E. Astrophys. J. Suppl. Ser. 116, 263 (1998) Composite m_{v} versus (V-I)_{0} diagram for template open clusters.

J/ApJS/117/135/  table1.dat  (64Kb)

Moffett T.J., Gieren W.P., Barnes T.G. III, Gomez M. Astrophys. J. Suppl. 117, 135 (1998) BV(RI)C Photometry of Cepheids in the Magellanic Clouds

J/ApJS/117/319/  notes.dat table1.dat table2.dat  (148Kb)

Appenzeller I., Thiering I., Zickgraf F.-J., Krautter J., Voges W., Chavarria C., Kneer R., Mujica R., Pakull M., Rosso C., Ruzicka F., Serrano A., Ziegler B. Astrophys. J. Suppl. Ser. 117, 319 (1998) Identification of a complete sample of northern ROSAT All-Sky Survey X-ray sources. III. The catalog.

J/ApJS/117/361/  notes.dat table1.dat  (123Kb)

Condon J.J., Kaplan D.L. Astrophys. J. Suppl. Ser. 117, 361 (1998) Planetary nebulae in the NRAO VLA sky survey

J/ApJS/117/427/  tab6-45.dat table3.dat table4.dat table5.dat table46.dat  (392Kb)

Nummelin A., Bergman P., Hjalmarson A., Friberg P., Irvine W.M., Millar T.J., Ohishi M., Saito S. Astrophys. J. Suppl. Ser. 117, 427 (1998) A three-position spectral line survey of Sagittarius B2 between 218 and 263 GHz. I. The observational data.

J/ApJS/117/587/  table1.dat table2.dat  (84Kb)

Heintz W.D. Astrophys. J. Suppl. Ser. 117, 587 (1998) Observations of double stars. XVIII.

J/ApJS/118/177/  table2.dat  (60Kb)

Kodaira K., Miyazaki S., Vansevicius V., Tamura M., Tokunaga A., Kobayashi N. Astrophys. J. Suppl. Ser. 118, 177 (1998) Near-infrared stellar photometry of the M31 spiral arm around OB association A24

J/ApJS/118/201/  table3.dat table4.dat  (36Kb)

Lazio T.J.W., Cordes J.M. Astrophys. J. Suppl. Ser. 118, 201 (1998) Hyperstrong radio-wave scattering in the Galactic Center. I. A survey for extragalactic sources seen through the Galactic Center.

J/ApJS/119/83/  notes.dat refs.dat table1.dat table2.dat table4.dat  (96Kb)

Parsons S.B., Ake T.B. Astrophys. J. Suppl. Ser. 119, 83 (1998) Ultraviolet and optical studies of binaries with luminous cool primaries and hot companions. V. The entire IUE sample.

J/ApJS/119/141/  table4.dat table5.dat table6.dat table9.dat  (65Kb)

Sloan G.C., Price S.D. Astrophys. J. Suppl. Ser. 119, 141 (1998) The infrared spectral classification of oxygen-rich dust shells.

J/ApJS/119/159/  table1.dat table3.dat table4.dat  (39Kb)

Spitzak J.G., Schneider S.E. Astrophys. J. Suppl. Ser., 119, 159-180 (1998) A deep survey of H I-selected Galaxies: the sample and the data.

J/ApJS/119/189/ field01h.dat field10h.dat field14h.dat field17h.dat field21h.dat field22h.dat fig1.ps fig2.ps fig3.ps table1.dat table1.ps table2.dat table2.ps table3.dat table3.ps (4.5 Mb) Osmer P.S., Kennefick, J.D., Hall P.B., Green R.F. Astrophys. J. Suppl. Ser. 119, 189 (1998) A Deep Multicolor Survey. IV. The Electronic Stellar Catalog

J/ApJS/119/207/ refs.dat sp.dat sp/* table1.dat table2.dat table4.dat table4b.dat table5.dat table5b.dat table6.dat table7.dat table7b.dat (8.9 Mb) Holberg J.B., Barstow M.A., Sion E.M. Astrophys. J. Suppl. Ser., 119, 207-238 (1998) A high-dispersion spectroscopic survey of the hot white dwarfs: the IUE NEWSIPS SWP echelle data set.

J/ApJS/119/277/  refs.dat table1.dat table3.dat  (88Kb)

Grogin N.A., Geller M.J., Huchra J.P. Astrophys. J. Suppl. Ser. 119, 277 (1998) A complete redshift survey to the Zwicky catalog limit in a 2^{h}x15^{∘}region around 3C 273.

J/ApJS/120/113/  table2.dat  (23Kb)

Esteban C., Peimbert M., Torres-Peimbert S., Garcia-Rojas J., Rodriguez M. Astrophys. J. Suppl. Ser. 120, 113 (1999) Faint Emission Lines and Temperature Fluctuations in M8

J/ApJS/120/147/  tab3-16.dat table1.dat table2.dat  (21Kb)

Marvel K.B., Shukla H., Rhee G. Astrophys. J. Suppl. Ser. 120, 147 (1999) A VLA search for head-tail radio sources in Abell clusters.

J/ApJS/120/247/  table1.dat table4.dat  (23Kb)

Gordon S.M., Duric N., Kirshner R.P., Goss W.M., Viallefond F. Astrophys. J. Suppl. Ser. 120, 247 (1999) A new sample of radio-selected and optically confirmed supernova remnants in M33

J/ApJS/120/265/  table5.dat  (589Kb)

Ibata R.A., Richer H.B., Fahlman G.G., Bolte M., Bond H.E., Hesser J.E., Pryor C., Stetson P.B. Astrophys. J. Suppl. Ser. 120, 265 (1999) Hubble Space Telescope photometry of the globular cluster M4.

J/ApJS/120/399/  table2.dat  (21Kb)

Hurley K., Briggs M. S., Kippen R. M., Kouveliotou C., Meegan C., Fishman G., Cline T., Boer M. Astrophys. J. Suppl. Ser. 120, 399 (1999) The ULYSSES Supplement to the BATSE 3B Catalog of Cosmic Gamma-Ray Bursts

J/ApJS/120/409/  table1.dat  (39Kb)

Vestrand W. T., Share G. H., Murphy R. J., Forrest D. J., Rieger E., Chupp E. L., Kanbach G. Astrophys. J. Suppl. Ser. 120, 409 (1999) The Solar Maximum Mission Atlas of Gamma-Ray Flares

J/ApJS/121/1 See III/210/

J/ApJS/121/171/  table1.dat  (12Kb)

Weisberg J. M., Cordes J. M., Lundgren S. C., Dawson B. R., Despotes J. T., Morgan J. J., Weitz K. A., ZinK E. C., Backer D. C. Astrophys. J. Suppl. Ser. 121, 171 (1999) Arecibo 1418 MHz Polarimetry of 98 Pulsars: Full Stokes Profiles and Morphological Classifications

J/ApJS/121/265/  table1.dat table2.dat  (96Kb)

Deb N.C., Aggarwal K.M., Msezane A.Z. Astrophys. J. Suppl. Ser. 121, 265 (1999) Oscillator Strengths and Radiative Rates for Transitions in Fe XV

J/ApJS/121/287/  cfa2s.dat clusters.dat notes.dat refs.dat  (341Kb)

Huchra J.P., Vogeley M.S., Geller M.J. Astrophys. J. Suppl. Ser. 121, 287 (1999)

The CfA redshift survey: data for the South Galactic Cap.

J/ApJS/121/473/  ref.dat table1.dat table2.dat  (34Kb)

Bassani L., Dadina M., Maiolino R., Salvati M., Risaliti G., Della Ceca R., Matt G., Zamorani G. Astrophys. J. Suppl. Ser. 121, 473 (1999) A Three-dimensional Diagnostic Diagram for Seyfert 2 Galaxies: Probing X-Ray Absorption and Compton Thickness

J/ApJS/121/591/  table1.dat  (32Kb)

Coutinho L.H., Trigueiros A.G. Astrophys. J. Suppl. Ser. 121, 591 (1999) Weighted Oscillator Strengths and Lifetimes for the Si VI Spectrum

J/ApJS/122/51/  table2.dat table4a.dat table4b.dat  (141Kb)

Dressler A., Smail I., Poggianti B.M., Butcher H., Couch W.J., Ellis R.S., Oemler A.Jr Astrophys. J. Suppl. Ser. 122, 51 (1999) A spectroscopic catalog of 10 distant rich clusters of galaxies.

J/ApJS/122/81/  refs.dat table2.dat table3.dat  (21Kb)

Martel, A. R., Baum S. A., Sparks W. B., Wyckoff E., Biretta J. A., Golombek D., Macchetto F. D., de Koff S., McCarthy P. J., Miley G. K. Astrophys. J. Suppl. Ser. 122, 81 (1999) Hubble Space Telescope Snapshot Survey of 3CR Radio Source Counterparts. III. Radio Galaxies with z 0.1

J/ApJS/122/109/  table2.dat  (11Kb)

Takamiya M. Astrophys. J. Suppl. Ser. 122, 109 (1999) Galaxy structural parameters: star formation rate and evolution with redshift.

J/ApJS/122/221/  refs.dat table1.dat table3.dat  (32Kb)

Meixner M., Ueta T., Dayal A., Hora J. L., Fazio G. Hrivnak B. J., Skinner C. J., Hoffmann W. F., Deutsch L. K. Astrophys. J. Suppl. Ser. 122, 221 (1999) A Mid-infrared Imaging Survey of Proto-Planetary Nebula Candidates

J/ApJS/122/355/ table4.dat table5.dat table6.dat table7.dat table8.dat table9.dat table10.dat table11.dat table12.dat table13.dat table14.dat table15.dat table16.dat (93Kb) Vanden Berk D. E., Lauroesch J. T., Stoughton C., Szalay A. S., Koo D. C., Crotts A. P. S., Blades J. C., Melott A. L., Boyle B. J., Broadhurst T. J., York D. G. Astrophys. J. Suppl. Ser. 122, 355 (1999) Clustering Properties of Low-Redshift QSO Absorption Systems Toward the Galactic Poles

J/ApJS/122/415/  table2.dat  (14Kb)

Alonso O., Garcia-Dabo C.E., Zamorano J., Gallego J., Rego M. Astrophys. J. Suppl. Ser. 122, 415 (1999) Survey for emission-line galaxies: Universidad Complutense de Madrid list 3.

J/ApJS/122/431/  table2.dat table3.dat  (19Kb)

Chandar R., Bianchi L., Ford H. C. Astrophys. J. Suppl. Ser. 122, 431 (1999) Star Clusters in M33. I. Detection, Morphology, and Photometry

J/ApJS/122/465 See IX/20A/

J/ApJS/123/41/  table2.dat  (148Kb)

Machalski J., Condon J.J. Astrophys. J. Suppl. Ser. 123, 41 (1999) Radio emission from galaxies in the Las Campanas redshift survey.

J/ApJS/123/79/  3eg.dat fluxes.dat notes.dat refs.dat table1.dat  (327Kb)

Hartman R.C., Bertsch D.L., Bloom S.D., Chen A.W., Deines-Jones P., Esposito J.A., Fichtel C.E., Friedlander D.P., Hunter S.D., McDonald L.M., Sreekumar P., Thompson D.J., Jones B.B., Lin Y.C., Michelson P.F., Nolan P.L., Tompkins W.F., Kanbach G., Mayer-Hasselwander H.A., Mucke A., Pohl M., Reimer O., Kniffen D.A., Schneid E.J., Von Montigny C., Mukherjee R., Dingus B.L. Astrophys. J. Suppl. Ser. 123, 79 (1999) The third EGRET catalog of high-energy gamma-ray sources.

J/ApJS/123/219/  notes.dat table1.dat table2.dat table3.dat table4.dat  (49Kb)

Condon J.J., Kaplan D.L., Terzian Y. Astrophys. J. Suppl. Ser. 123, 219 (1999) Infrared planetary nebulae in the NRAO VLA Sky Survey.

J/ApJS/123/233/  refs.dat table1.dat table2.dat table3.dat  (66Kb)

Lee C.W., Myers P.C. Astrophys. J. Suppl. Ser. 123, 233 (1999) A catalog of optically selected cores.

J/ApJS/123/283/  table3.dat  (28Kb)

Montes D., Ramsey L.W., Welty A.D. Astrophys. J. Suppl. Ser. 123, 283 (1999) Library of Medium-Resolution Fiber Optic Echelle Spectra of F, G, K, and M Field Dwarfs to Giant Stars

J/ApJS/123/295/  table1.dat table2.dat table3.dat  (75Kb)

Gupta G.P., Tayal S.S. Astrophys. J. Suppl. Ser. 123, 295 (1999) Electron Impact Excitation of Fine-Structure Levels in Sulfur-like Fe XI

J/ApJS/123/311/  table1.dat table2.dat table4a.dat table4b.dat  (203Kb)

Aggarwal K.M., Keenan F.P. Astrophys. J. Suppl. Ser. 123, 311 (1999) Excitation Rate Coefficients for Fine-Structure Transitions in O III

J/ApJS/123/377/  table5.dat  (301Kb)

Newberg H.J., Richards G.T., Richmond M., Fan X. Astrophys. J. Suppl. Ser. 123, 377 (1999) Catalog of four-color photometry of stars, galaxies, and QSOs using SDSS filters

J/ApJS/123/639/  table2.dat table3.dat table4.dat  (154Kb)

Norris J.E., Ryan S.G., Beers T.C. Astrophys. J. Suppl. Ser. 123, 639 (1999) A search for stars of very low metal abundance. III. UBV photometry of metal-weak candidates.

J/ApJS/124/127/  tablea1.dat tablea2.dat  (88Kb)

Pahre M. A. Astrophys. J. Suppl. Ser. 124, 127 (1999) Near-infrared Imaging of Early-Type Galaxies. II. Global Photometric Parameters.

J/ApJS/124/195/  fits/* fluxes.dat objects.dat sp/* table1.dat table8.dat  (3.3 Mb)

Hora J.L., Latter W.B., Deutsch L.K. Astrophys. J. Suppl. Ser., 124, 195-240 (1999) Investigating the near-infrared properties of planetary nebulae. II. Medium-resolution spectra.

J/ApJS/124/285/  table1.dat  (28Kb)

Reid R. I., Kronberg P. P., Perley R. A. Astrophys. J. Suppl. Ser. 124, 285 (1999) VLA Images at 5 GHz of 212 Southern Extragalactic Objects

J/ApJS/124/527/  table2.dat table3.dat  (165Kb)

Melendez J., Barbuy B. Astrophys. J. Suppl. Ser. 124, 527 (1999) Oscillator Strengths and Damping Constants for Atomic Lines in the J and H Bands

J/ApJS/125/161/  refs.dat tablea1.dat tablea2.dat tablea3.dat tablea4.dat  (123Kb)

Jijina J., Myers P.C., Adams F.C. Astrophys. J. Suppl. Ser. 125, 161 (1999) Dense cores mapped in ammonia: a database.

J/ApJS/125/349/  table2.dat  (34Kb)

Hunt L. K., Malkan M. A., Rush B., Bicay M. D., Nelson B. O., Stanga R. M., Webb W. Astrophys. J. Suppl. Ser. 125, 349 (1999) Morphology of the 12 Micron Seyfert Galaxies. II. Optical and Near-Infrared Image Atlas

J/ApJS/125/409/  table1.dat  (1.2 Mb)

Cotton W.D., Condon J.J. Astrophys. J. Suppl. Ser. 125, 409 (1999) Arcsecond Positions of UGC Galaxies

J/ApJS/125/427/  table1.dat table2.dat  (10Kb)

Rodriguez L.F., Anglada G., Curiel S. Astrophys. J. Suppl. Ser. 125, 427 (1999) The nature of the radio continuum sources embedded in the HH 7-11 region and its surroundings.

J/ApJS/126/1/  table6.dat  (39Kb)

Quintana H., Infante L., Fouque P., Carrasco E.R., Cuevas H., Hertling G., Nunez I.

Astrophys. J. Suppl. Ser. 126, 1 (2000) Photometry of the central regions in a complete sample of high Bautz-Morgan-Type Abell clusters. I. Observations and data calibration.

J/ApJS/126/19/  table1.dat  (16Kb)

Preece R.D., Briggs M.S., Mallozzi R.S., Pendleton G.N., Paciesas W.S., Band D.L. Astrophys. J. Suppl. Ser. 126, 19 (2000) The BATSE gamma-ray burst spectral catalog. I. High time resolution spectroscopy of bright bursts using high energy resolution data.

J/ApJS/126/85/  table3.dat table8.dat table9.dat table10.dat  (17Kb)

Kerton C.R., Martin P.G. Astrophys. J. Suppl. Ser. 126, 85 (2000) A mid-infrared galaxy atlas (MIGA).

J/ApJS/126/133/ fits/* ref.dat table2.dat table3.dat table4.dat table5.dat table6.dat table7.dat table8.dat (6.0 Mb) White R.L., Becker R.H., Gregg M.D., Laurent-Muehleisen S.A., Brotherton M.S., Impey C.D., Petry C.E., Foltz C.B., Chaffee F.H., Richards G.T., Oegerle W.R., Helfand D.J., McMahon R.G., Cabanela J.E. Astrophys. J. Suppl. Ser. 126, 133 (2000) The FIRST bright quasar survey. II. 60 nights and 1200 spectra later

J/ApJS/126/209/ notes.dat refs.dat table1.dat table2.dat table3.dat table4.dat table5.dat table6.dat (93Kb) Romer A.K., Nichol R.C., Holden B.P., Ulmer M.P., Pildis R.A., Merrelli A.J., Adami C., Burke D.J., Collins C.A., Metevier A.J., Kron R.G., Commons K. Astrophys. J. Suppl. Ser. 126, 209 (2000) The bright SHARC survey: the cluster catalog.

J/ApJS/126/271/  table1.dat table3.dat  (61Kb)

Jansen R.A., Franx M., Fabricant D., Caldwell N. Astrophys. J. Suppl. Ser. 126, 271 (2000) Surface photometry of nearby field galaxies: the data.

J/ApJS/126/331/  notes.dat table1.dat table2.dat table4.dat table5.dat table6.dat  (85Kb)

Jansen R.A., Fabricant D., Franx M., Caldwell N. Astrophys. J. Suppl. Ser. 126, 331 (2000) Spectrophotometry of nearby field galaxies: the data.

J/ApJS/126/461/  table2.dat  (25Kb)

Zhao G., Qiu H.M., Chen Y.Q., Li Z.W. Astrophys. J. Suppl. 126, 461 (2000) Equivalent Widths in the Spectrum of Sirius

J/ApJS/127/1/  table2.dat table3.dat table4.dat  (370Kb)

Hogg D.W., Pahre M.A., Adelberger K.L., Blandford R., Cohen J.G., Gautier T.N., Jarrett T., Neugebauer G., Steidel C.C. Astrophys. J. Suppl. Ser. 127, 1 (2000) Caltech Faint Galaxy Redshift Survey. IX. Source Detection and Photometry in the Hubble Deep Field Region

J/ApJS/127/27 See VII/209A/

J/ApJS/127/39/  galaxies.dat gaseous.dat stellar.dat  (102Kb)

Caon C., Macchetto D., Pastoriza M. Astrophys. J. Suppl. Ser. 127, 39 (2000) A survey of the ISM in early-type galaxies. III. Stellar and gas kinematics

J/ApJS/127/79/  table3.dat  (110Kb)

Ling J.C., Wheaton W. A., Wallyn P., Skelton R.T., Mahoney W.A., Radocinski R.G., Callas J.L., Ling N.F., Tumer E., Shubert R. Astrophys. J. Suppl. Ser. 127, 79 (2000) A BATSE Earth occultation catalog of 0.03-1.8 MeV Gamma-Ray source spectra and light curves for phases 1-3 (1999-1994)

J/ApJS/128/17/  table1.dat  (30Kb)

Fey A.L., Charlot P. Astrophys. J. Suppl. Ser. 128, 17 (2000) VLBA observations of radio reference frame sources. III. Astrometric suitability of an additional 225 sources.

J/ApJS/128/99/  table2.dat table3.dat table4.dat  (30Kb)

O’Neil K., Bothun G.D., Impey C.D.

Astrophys. J. Suppl. Ser. 128, 99 (2000) Structural characteristics of faint galaxies serendipitously discovered with the Hubble space telescope WFPC2.

J/ApJS/129/435/ refs.dat table1.dat table2.dat table3.dat table4.dat table5.dat table6.dat table7.dat table8.dat table9.dat (116Kb) Bohringer H., Voges W., Huchra J.P., McLean B., Giacconi R., Rosati P., Burg R., Mader J., Schuecker P., Simic D., Komossa S., Reiprich T.H., Retzlaff J., Trumper J. Astrophys. J. Suppl. Ser. 129, 435 (2000) The northern ROSAT all-sky (NORAS) galaxy cluster survey. I. X-ray properties of clusters detected as extended X-ray sources.

J/ApJS/129/475/  cnoc0223.dat  (6.3 Mb)

Yee H.K.C., Morris S.L., Lin H., Carlberg R.G., Hall P.B., Sawicki M., Patton D.R., Wirth G.D., Ellingson E., Shepherd C.W. Astrophys. J. Suppl. Ser. 129, 475 (2000) The CNOC2 field galaxy redshift survey. I. The survey and the catalog for the patch CNOC 0223+00.

J/ApJS/129/547/  notes.dat table1.dat table2.dat  (234Kb)

Bauer F.E., Condon J.J., Thuan T.X., Broderick J.J. Astrophys. J. Suppl. Ser. 129, 547 (2000) RBSC-NVSS sample. I. Radio and optical identifications of a complete sample of 1556 bright X-ray sources.

J/ApJS/130/177/  table2.dat table3.dat table4.dat  (56Kb)

Rosenberg J.L., Schneider S.E. Astrophys. J. Suppl. Ser. 130, 177 (2000) The Arecibo Dual-Beam Survey: Arecibo and VLA observations.

J/ApJS/130/237/  table1.dat table3.dat table7.dat  (861Kb)

Tucker D.L., Oemler A.Jr, Hashimoto Y., Shectman S.A., Kirshner R.P., Lin H., Landy S.D., Schechter P.L., Allam S.S. Astrophys. J. Suppl. Ser. 130, 237 (2000) Loose groups of galaxies in the Las Campanas Redshift Survey.

J/ApJS/131/47/  table1.dat table2.dat  (37Kb)

Eder J.A., Schombert J.M. Astrophys. J. Suppl. Ser. 131, 47 (2000) Gas-rich dwarfs from the PSS-II. III. H I profiles and dynamical masses.

J/ApJS/131/95/  table1.dat table2.dat  (103Kb)

Fomalont E.B., Frey S., Paragi Z., Gurvits L.I., Scott W.K., Taylor A.R., Edwards P.G., Hirabayashi H. Astrophys. J. Suppl. Ser. 131, 95 (2000) The VSOP 5 GHz continuum survey: the prelaunch VLBA observations.

J/ApJS/131/185/  table1.dat  (20Kb)

Stanford S.A., Stern D., Van Breugel W., De Breuck C. Astrophys. J. Suppl. Ser. 131, 185 (2000) The FIRST sample of ultraluminous infrared galaxies at high redshift. I. Sample and near-infrared morphologies.

J/ApJS/131/561/  table1.dat  (22Kb)

Scardia M., Prieur J.-L., Aristidi E., Koechlin L. Astrophys. J. Suppl. Ser. 131, 561 (2000) Speckle observations of double stars with PISCO at Pic du midi: measurements in 1998.

J/ApJS/132/37/  table1.dat table2.dat table3.dat table4.dat  (111Kb)

Kewley L.J., Heisler C.A., Dopita M.A., Lumsden S. Astrophys. J. Suppl. Ser. 132, 37 (2001) Optical classification of southern warm infrared galaxies.

J/ApJS/133/1/  table1.dat table2.dat  (230Kb)

Ueda Y., Ishisaki Y., Takahashi T., Makishima K., Ohashi T. Astrophys. J. Suppl. Ser. 133, 1 (2001) The ASCA Medium Sensitivity Survey (the GIS catalog project): source catalog.

J/ApJS/133/77/  table1.dat table3.dat table4.dat table5.dat  (48Kb)

Ho L.C., Ulvestad J.S. Astrophys. J. Suppl. Ser. 133, 77 (2001) Radio continuum survey of an optically selected sample of nearby Seyfert galaxies.

J/ApJS/133/119/  table1.dat table2.dat  (46Kb)

Ruiz M.T., Wischnjewsky M., Rojo P.M., Gonzalez L.E.

Astrophys. J. Suppl. Ser. 133, 119 (2001) Calan-ESO proper-motion catalog.

J/ApJS/133/413/  table1.dat table2.dat  (35Kb)

Bergeron P., Leggett S.K., Ruiz M.T. Astrophys. J. Suppl. Ser. 133, 413 (2001) Photometric and spectroscopic analysis of cool white dwarfs with trigonometric parallax measurements.

J/ApJS/134/35/  table1.dat table5.dat table6.dat  (3.2 Mb)

Forster K., Green P.J., Aldcroft T.L., Vestergaard M., Foltz C.B., Hewett P.C. Astrophys. J. Suppl. Ser. 134, 35 (2001) Emission line properties of the large bright quasar survey.

J/ApJS/134/77/  table2.dat table5.dat  (32Kb)

Sugizaki M., Mitsuda K., Kaneda H., Matsuzaki K., Yamauchi S., Koyama K. Astrophys. J. Suppl. Ser. 134, 77 (2001) Faint X-ray sources resolved in the ASCA galactic plane survey and their contribution to the galactic ridge X-ray emission.

J/ApJS/134/103/  table2.dat  (27Kb)

Barrado y Navascues D., Stauffer J.R., Briceno C., Patten B., Hambly N.C., Adams J.D. Astrophys. J. Suppl. Ser. 134, 103 (2001) Very low-mass stars and brown dwarfs of the young open cluster IC 2391.

J/ApJS/134/355/  table5.dat table6.dat  (59Kb)

Miller N.A., Owen F.N. Astrophys. J. Suppl. Ser. 134, 355 (2001) The radio galaxy populations of nearby Northern Abell clusters

J/ApJS/134/385/  table5.dat table6.dat table8.dat  (146Kb)

Kommers J.M., Lewin W.H.G., Kouveliotou C., Van Paradijs J., Pendleton G.N., Meegan C.A., Fishman G.J. Astrophys. J. Suppl. Ser. 134, 385 (2001) A non-triggered burst supplement to the BATSE gamma-ray burst catalogs.

J/ApJS/135/41/  table1.dat table2.dat table3.dat table6.dat  (25Kb)

Fernandez-Soto A., Lanzetta K.M., Chen H.-W., Pascarelle S.M., Yahata N. Astrophys. J. Suppl. Ser. 135, 41 (2001) On the compared accuracy and reliability of spectroscopic and photometric redshift measurements.

J/ApJS/135/155/  table1.dat table2.dat table3.dat table4.dat  (90Kb)

Mattox J.R., Hartman R.C., Reimer O. Astrophys. J. Suppl. Ser. 135, 155 (2001) A quantitative evaluation of potential radio identifications for 3EG EGRET sources.

J/ApJS/135/227/ table1.dat table2.dat table3.dat table4.dat table5.dat table6.dat table7.dat table8.dat (72Kb) Becker R.H., White R.L., Gregg M.D., Laurent-Muehleisen S.A., Brotherton M.S., Impey C.D., Chaffee F.H., Richards G.T., Helfand D.J., Lacy M., Courbin F., Proctor D.D. Astrophys. J. Suppl. Ser. 135, 227 (2001) The FIRST bright quasar survey. III. The South Galactic Cap.

J/ApJS/136/265/  table1.dat  (20Kb)

Lazio T.J.W., Waltman E.B., Ghigo F.D., Fiedler R.L., Foster R.S., Johnston K.J. Astrophys. J. Suppl. Ser. 136, 265 (2001) A dual-frequency, multiyear monitoring program of compact radio sources.

J/ApJS/136/463/  cloud.dat notes.dat refs.dat table2a.dat table2b.dat telescop.dat  (284Kb)

Wakker B.P. Astrophys. J. Suppl. Ser. 136, 463 (2001) Distances and metallicities of high- and intermediate-velocity clouds.

J/ApJS/136/631/  refs.dat table1.dat table3.dat table4.dat  (120Kb)

Savage B.D., Meade M.R., Sembach K.R. Astrophys. J. Suppl. Ser. 136, 631 (2001) IUE absorption-line observations of the moderately and highly ionized interstellar medium toward 164 earlytype stars.

J/ApJS/137/117/  table1.dat table2.dat  (85Kb)

Gonzalez A.H., Zaritsky D., Dalcanton J.J., Nelson A. Astrophys. J. Suppl. Ser. 137, 117 (2001) The Las Campanas Distant Cluster Survey: the catalog.

J/ApJS/137/209/  stars.dat table1.dat  (117Kb)

Berdnikov L.N., Turner D.G. Astrophys. J. Suppl. Ser. 137, 209 (2001) Photoelectric observations of southern Cepheids in 2001.

J/ApJS/137/279/  table3.dat  (48Kb)

Mobasher B., Bridges T.J., Carter D., Poggianti B.M., Komiyama Y., Kashikawa N., Doi M., Iye M., Okamura S., Sekiguchi M., Shimasaku K., Yagi M., Yasuda N. Astrophys. J. Suppl. Ser. 137, 279 (2001) A photometric and spectroscopic study of dwarf and giant galaxies in the Coma cluster. II. Spectroscopic observations.

J/ApJS/139/369/  table2.dat table3.dat table4.dat table5.dat  (78Kb)

Giacconi R., Zirm A., Wang J., Rosati P., Nonino M., Tozzi P., Gilli R., Mainieri V., Hasinger G., Kewley L., Bergeron J., Borgani S., Gilmozzi R., Grogin N., Koekemoer A., Schreier E., Zheng W., Norman C. Astrophys. J. Suppl. Ser. 139, 369 (2002) Chandra Deep Field South: the 1 Ms catalog.

J/ApJS/140/143/  spectra/* table1.dat table2.dat table3.dat table4.dat  (39.3 Mb)

Bechtold J., Dobrzycki A., Wilden B., Morita M., Scott J., Dobrzycka D., Tran K.-V., Aldcroft T.L. Astrophys. J. Suppl. Ser. 140, 143 (2002) A uniform analysis of the Lyαforest at z=0-5. III. Hubble Space Telescope Faint Object Spectrograph spectral atlas.

J/ApJS/140/239/  refs.dat table3.dat  (35Kb)

Cruddace R., Voges W., Bohringer H., Collins C.A., Romer A.K., MacGillivray H., Yentis D., Schuecker P., Ebeling H., de Grandi S. Astrophys. J. Suppl. Ser. 140, 239 (2002) The ROSAT All-Sky Survey: a catalog of clusters of galaxies in a region of 1 steradian around the South Galactic Pole.

J/ApJS/140/265/  table2.dat table3.dat table4.dat table5.dat table6.dat  (40Kb)

Perlman E.S., Horner D.J., Jones L.R., Scharf C.A., Ebeling H., Wegner G., Malkan M. Astrophys. J. Suppl. Ser. 140, 265 (2002) The WARPS survey. VI. Galaxy cluster and source identifications from phase I.

J/ApJS/140/331/  table1.dat table3.dat table4.dat  (98Kb)

Lockman F.J., Murphy E.M., Petty-Powell S., Urick V.J. Astrophys. J. Suppl. Ser., 140, 331-365 (2002) A very sensitive 21 centimeter survey for galactic high-velocity H I.

J/ApJS/141/13/  table2.dat table3.dat table4.dat  (147Kb)

Beasley A.J., Gordon D., Peck A.B., Petrov L., MacMillan D.S., Fomalont E.B., Ma C. Astrophys. J. Suppl. Ser. 141, 13 (2002) The VLBA calibrator Survey - VCS1.

J/ApJS/141/23/  table1.dat table2a.dat table2b.dat table2c.dat  (161Kb)

Enya K., Yoshii Y., Kobayashi Y., Minezaki T., Suganuma M., Tomita H., Peterson B.A. Astrophys. J. Suppl. Ser. 141, 23 (2002) JHK’ imaging photometry of Seyfert 1 active galactic nuclei and quasars. I. Multiaperture photometry.

J/ApJS/141/31/  table1.dat  (24Kb)

Enya K., Yoshii Y., Kobayashi Y., Minezaki T., Suganuma M., Tomita H., Peterson B.A. Astrophys. J. Suppl. Ser. 141, 31 (2002) JHK’ imaging photometry of Seyfert 1 active galactic nuclei and quasars. II. Observation of long-term variability.

J/ApJS/141/61/  table1.dat  (11Kb)

Gutierrez C.M., Azzaro M., Prada F. Astrophys. J. Suppl. Ser. 141, 61 (2002) The properties of satellite galaxies in external systems. I. Morphology and structural parameters.

J/ApJS/141/81 See II/236/

J/ApJS/141/157/  spectra/* table2.dat  (37.2 Mb)

Yang J., Jiang Z., Wang M., Ju B., Wang H. Astrophys. J. Suppl. Ser. 141, 157 (2002) A large-scale molecular line survey for cold IRAS sources in the Galaxy. I. The CO (J=1-0) data.

J/ApJS/141/187 See I/279/

J/ApJS/141/469/  table2.dat  (71Kb)

Pereyra A., Magalhaes A.M. Astrophys. J. Suppl. Ser. 141, 469 (2002) Polarimetry toward the IRAS Vela shell. I. The catalog.

J/ApJS/141/503/  table1.dat table2.dat table3.dat table4.dat  (62Kb)

Nidever D.L., Marcy G.W., Butler R.P., Fischer D.A., Vogt S.S. Astrophys. J. Suppl. Ser. 141, 503 (2002) Radial velocities of 889 late-type stars

J/ApJS/142/1/ table2a.dat table2b.dat table2c.dat table2d.dat table2e.dat table2f.dat table2g.dat table3a.dat table3d.dat table3e.dat table3f.dat table3g.dat table4.dat table5.dat (26.3 Mb) Simard L., Willmer C.N.A., Vogt N.P., Sarajedini V.L., Phillips A.C., Weiner B.J., Koo D.C., Im M., Illingworth G.D., Faber S.M. Astrophys. J. Suppl. Ser. 142, 1 (2002) The DEEP Groth Strip Survey. II. Hubble Space Telescope structural parameters of galaxies in the Groth Strip.

J/ApJS/142/71/  table3.dat table4.dat  (172Kb)

Sebo K.M., Rawson D., Mould J., Madore B.F., Putman M.E., Graham J.A., Freedman W.L., Gibson B.K., Germany L.M. Astrophys. J. Suppl. Ser. 142, 71 (2002) The Cepheid period-luminosity relation in the Large Magellanic Cloud.

J/ApJS/142/79/  table1.dat table3.dat  (24Kb)

Malkan M.A., Hicks E.K., Teplitz H.I., McLean I.M., Sugai H., Guichard J. Astrophys. J. Suppl. Ser. 142, 79 (2002) 1-1.4 micron spectral atlas of stars.

J/ApJS/142/153/ hjir-k.dat hjiv-k.dat hjrg-k.dat jirb-k.dat jiv-h.dat jr-h.dat jvrb-k.dat table1.dat table2.dat (886Kb) Stanford S.A., Eisenhardt P.R., Dickinson M., Holden B.P., de Propris R. Astrophys. J. Suppl. Ser. 142, 153 (2002) Optical and near-infrared photometry of distant galaxy clusters.

J/ApJS/143/1/  table4.dat  (63.1 Mb)

McMahon R.G., White R.L., Helfand D.J., Becker R.H. Astrophys. J. Suppl. Ser. 143, 1 (2002) Optical counterparts for 70,000 radio sources: APM identifications for the FIRST radio survey.

J/ApJS/143/25/  table1.dat table2.dat  (22Kb)

Colbert E.J.M., Ptak A.F. Astrophys. J. Suppl. Ser. 143, 25 (2002) A catalog of candidate intermediate-luminosity X-ray objects.

J/ApJS/143/73/  table1.dat  (20Kb)

Eskridge P.B., Frogel J.A., Pogge R.W., Quillen A.C., Berlind A.A., Davies R.L., Depoy D.L., Gilbert K.M., Houdashelt M.L., Kuchinski L.E., Ramirez S.V., Sellgren K., Stutz A., Terndrup D.M., Tiede G.P. Astrophys. J. Suppl. Ser. 143, 73 (2002) Near-infrared and optical morphology of spiral galaxies.

J/ApJS/143/257/  table1.dat table4.dat table5.dat  (1.0 Mb)

Kuraszkiewicz J.K., Green P.J., Forster K., Aldcroft T.L., Evans I.N., Koratkar A. Astrophys. J. Suppl. Ser. 143, 257 (2002) Emission line properties of active galactic nuclei from a pre-COSTAR Faint Object Spectrograph Hubble Space Telescope spectral atlas.

J/ApJS/143/277/  table1.dat table2.dat table3.dat  (55Kb)

Kim D.-C., Veilleux S., Sanders D.B. Astrophys. J. Suppl. Ser. 143, 277 (2002) Optical and near-infrared imaging of the IRAS 1Jy sample of ultraluminous galaxies. I. The atlas.

J/ApJS/143/315/  table1.dat table2.dat  (21Kb)

Veilleux S., Kim D.-C., Sanders D.B. Astrophys. J. Suppl. Ser. 143, 315 (2002) Optical and near-infrared imaging of the IRAS 1Jy sample of ultraluminous galaxies. II. The analysis.

J/ApJS/143/377/  table1.dat table4.dat table5.dat table6.dat table9.dat  (32Kb)

Wu W., Clayton G.C., Gordon K.D., Misselt K.A., Smith T.L., Calzetti D. Astrophys. J. Suppl. Ser. 143, 377 (2002) A new database of observed spectral energy distributions of nearby starburst galaxies from the ultraviolet to the far-infrared.

J/ApJS/143/513/  notes.dat table1.dat  (34Kb)

Ginestet N., Carquillat J.M. Astrophys. J. Suppl. Ser. 143, 513 (2002) Spectral classification of the hot components of a large sample of stars with composite spectra, and implication for the absolute magnitudes of the cool supergiant components.

J/ApJS/144/21/  refs.dat table1.dat table2.dat  (30Kb)

Robert C., Pellerin A., Aloisi A., Leitherer C., Hoopes C., Heckman T.M. Astrophys. J. Suppl. Ser. 144, 21 (2003) Synthetic high-resolution line spectra of star-forming galaxies below 1200Å.

J/ApJS/144/47/  table3.dat  (3.0 Mb)

Brunt C.M., Kerton C.R., Pomerleau C. Astrophys. J. Suppl. Ser. 144, 47 (2003) An outer galaxy molecular cloud catalog.

J/ApJS/144/213/  table2.dat table3.dat  (28Kb)

Swartz D.A., Ghosh K.K., McCollough M.L., Pannuti T.G., Tennant A.F., Wu K. Astrophys. J. Suppl. Ser. 144, 213 (2003) Chandra X-ray observations of the spiral galaxy M81.

J/ApJS/145/65/  table6.dat  (22Kb)

Feldmeier J.J., Ciardullo R., Jacoby G.H., Durrell P.R. Astrophys. J. Suppl. Ser. 145, 65 (2003) Intracluster planetary nebulae in the Virgo cluster. II. Imaging catalog.

J/ApJS/145/83/  table1.dat table2.dat table3.dat  (31Kb)

Winn J.N., Patnaik A.R., Wrobel J.M. Astrophys. J. Suppl. Ser. 145, 83 (2003) Interferometric phase calibration sources in the declination range 0^{∘}to -30^{∘}.

J/ApJS/145/199/  table1.dat table2.dat table3.dat table4.dat table5.dat  (149Kb)

Marziani P., Sulentic J.W., Zamanov R., Calvani M., Dultzin-Hacyan D., Bachev R., Zwitter T. Astrophys. J. Suppl. Ser. 145, 199 (2003) An optical spectroscopic atlas of low-redshift active galactic nuclei.

J/ApJS/145/213/  table1.dat  (26Kb)

Brown J.C., Taylor A.R., Jackel B.J. Astrophys. J. Suppl. Ser. 145, 213 (2003) Rotation measures of compact sources in the Canadian Galactic Plane Survey.

J/ApJS/145/245/  table3.dat table4.dat  (1.3 Mb)

Rave H.A., Zhao C., Newberg H.J., Yanny B., Schneider D.P., Brinkman J., Lamb D.Q. Astrophys. J. Suppl. Ser. 145, 245 (2003) Sloan Digital Sky Survey catalog of stars in the Draco dwarf spheroidal galaxy.

J/ApJS/145/329/  sources.dat table1.dat table2.dat  (19.3 Mb)

Heiles C., Troland T.H. Astrophys. J. Suppl. Ser. 145, 329 (2003) The Millennium Arecibo 21 centimeter Absorption-Line Survey. I. Techniques and gaussian fits.

J/ApJS/145/355/  table1.dat table2.dat table3.dat table4.dat tables.dat  (42Kb)

Genova R., Beckman J.E. Astrophys. J. Suppl. Ser. 145, 355 (2003) Kinematical structure of the local interstellar medium: the Galactic anticenter hemisphere.

J/ApJS/146/1/  notes.dat table1.dat table2.dat  (87Kb)

Wakker B.P., Savage B.D., Sembach K.R., Richter P., Meade M., Jenkins E.B., Shull J.M., Ake T.B., Blair W.P., Dixon W.V., Friedman S.D., Green J.C., Green R.F., Kruk J.W., Moos H.W., Murphy E.M., Oegerle W.R., Sahnow D.J., Sonneborn G., Wilkinson E., York D.G. Astrophys. J. Suppl. Ser. 146, 1 (2003) The Far Ultraviolet Spectroscopic Explorer Survey of O VI absorption in and near the galaxy.

J/ApJS/146/125/  table2.dat table7.dat  (31Kb)

Savage B.D., Sembach K.R., Wakker B.P., Richter P., Meade M., Jenkins E.B., Shull J.M., Moos H.W., Sonneborn G. Astrophys. J. Suppl. Ser. 146, 125 (2003) Distribution and kinematics of O VI in the Galactic halo.

J/ApJS/146/209/  table2.dat table3.dat table4.dat  (44Kb)

Teplitz H.I., Collins N.R., Gardner J.P., Hill R.S., Heap S.R., Lindler D.J., Rhodes J., Woodgate B.E.

Astrophys. J. Suppl. Ser. 146, 209 (2003) Emission-line galaxies in the Space Telescope Imaging Spectrograph Parallel Survey. I. Observations and data analysis.

J/ApJS/146/267/  table3.dat table4.dat  (27Kb)

Morrison G.E., Owen F.N., Ledlow M.J., Keel W.C., Hill J.M., Voges W., Herter T. Astrophys. J. Suppl. Ser. 146, 267 (2003) Radio-selected galaxies in very rich clusters at z =0.25. I. Multiwavelength observations and data reduction techniques.

J/ApJS/146/353/  refs.dat table1.dat table2.dat table3.dat table4.dat  (35Kb)

Martini P., Regan M.W., Mulchaey J.S., Pogge R.W. Astrophys. J. Suppl. Ser. 146, 353 (2003) Circumnuclear dust in nearby active and inactive galaxies. I. Data.

J/ApJS/147/1/  table1.dat table2.dat table3.dat table4.dat table5.dat  (25Kb)

Conselice C.J. Astrophys. J. Suppl. Ser. 147, 1 (2003) The relationship between stellar light distributions of galaxies and their formation histories.

J/ApJS/147/29/  table1.dat table3.dat table5.dat  (39Kb)

Gil De Paz A., Madore B.F., Pevunova O. Astrophys. J. Suppl. Ser. 147, 29 (2003) Palomar/Las Campanas imaging atlas of blue compact dwarf galaxies. I. Images and integrated photometry.

J/ApJS/147/61/  refs.dat table9.dat  (17Kb)

Welty D.E., Hobbs L.M., Morton D.C. Astrophys. J. Suppl. Ser. 147, 61 (2003) High-resolution observations of interstellar Ca I absorption. Implications for depletions and electron densities in diffuse clouds.

J/ApJS/147/103/  table1.dat  (20Kb)

Griffin R.F., Suchkov A.A. Astrophys. J. Suppl. Ser. 147, 103 (2003) The nature of overluminous F stars observed in a radial-velocity survey.

J/ApJS/147/145/  table2.dat  (23Kb)

Holberg J.B., Barstow M.A., Burleigh M.R. Astrophys. J. Suppl. Ser. 147, 145 (2003) An archive of IUE low-dispersion spectra of the white dwarf stars.

J/ApJS/147/305/  refs.dat table1.dat table2.dat table3.dat table6.dat table7.dat table9.dat  (67Kb)

Valenti J.A., Fallon A.A., Johns-Krull C.M. Astrophys. J. Suppl. Ser. 147, 305 (2003) An IUE atlas of pre-main-sequence stars. III. Co-added final archive spectra from the long-wavelength cameras.

J/ApJS/148/97/  wmap1.dat  (28Kb)

Bennett C.L., Hill R.S., Hinshaw G., Nolta M.R., Odegard N., Page L., Spergel D.N., Weiland J.L., Wright E.L., Halpern M., Jarosik N., Kogut A., Limon M., Meyer S.S., Tucker G.S., Wollack E. Astrophys. J. Suppl. Ser. 148, 97 (2003) First-year Wilkinson microwave anisotropy probe (WMAP) first year observations: foreground emission.

J/ApJS/148/243/  catalog.dat  (63Kb)

Bahcall N.A., McKay T.A., Annis J., Kim R.S.J., Dong F., Hansen S., Goto T., Gunn J.E., Miller C., Nichol R.C., Postman M., Schneider D., Schroeder J., Voges W., Brinkmann J., Fukugita M. Astrophys. J. Suppl. Ser. 148, 243 (2003) A merged catalog of clusters of galaxies from Early Sloan Digital Sky Survey Data.

J/ApJS/148/275/  table1.dat table2.dat table3.dat table4.dat  (55Kb)

Akiyama M., Ueda Y., Ohta K., Takahashi T., Yamada T. Astrophys. J. Suppl. Ser. 148, 275 (2003) Optical identification of the ASCA medium sensitivity survey in the northern sky: nature of hard X-Ray-Selected luminous active galactic nuclei.

J/ApJS/149/29/  notes.dat table3.dat  (92Kb)

Gioia I.M., Henry J.P., Mullis C.R., Bohringer H., Briel U.G., Voges W., Huchra J.P. Astrophys. J. Suppl. Ser. 149, 29 (2003) The ROSAT North Ecliptic Pole Survey: the optical identifications.

J/ApJS/149/53/  table2.dat table4.dat  (94Kb)

Yuan Q., Zhou X., Jiang Z.

Astrophys. J. Suppl. Ser. 149, 53 (2003) Multicolor photometry of the galaxies in A2255 by the Beijing-Arizona-Taiwan-Connecticut Survey and Sloan Digital Sky Survey.

J/ApJS/149/123/  table2.dat table7.dat  (48Kb)

Setia Gunawan D.Y.A., de Bruyn A.G., van der Hucht K.A., Williams P.M. Astrophys. J. Suppl. Ser. 149, 123 (2003) A Westerbork Synthesis Radio Telescope 1400 and 350 MHz continuum survey of the Cygnus OB2 association, in search of hot massive stars.

J/ApJS/149/157/  table3.dat  (148Kb)

Sharpee B., Williams R., Baldwin J.A., van Hoof P.A.M. Astrophys. J. Suppl. Ser. 149, 157 (2003) Introducing EMILI: computer-aided emission line identification.

J/ApJS/149/327/  table1.dat table2.dat  (48Kb)

Grauer A.D., Rieke M.J., Quillen A.C. Astrophys. J. Suppl. Ser., 149, 327-342 (2003) Infrared observations of galaxies in the local universe. II. 391 calibrated images with photometric and structural measurements.

J/ApJS/149/343/  table1.dat  (16Kb)

Engargiola G., Plambeck R.L., Rosolowsky E., Blitz L. Astrophys. J. Suppl. Ser., 149, 343 (2003) Giant molecular clouds in M33. I. BIMA all-disk survey.

J/ApJS/149/365/  table1.dat table2.dat table3.dat  (21Kb)

Jackson T., Werner M., Gautier Iii T.N. Astrophys. J. Suppl. Ser., 149, 365-373 (2003) A catalog of bright filamentary structures in the interstellar medium.

J/ApJS/149/405 See II/249/

J/ApJS/150/1/  table4a.dat table4b.dat table5a.dat table5b.dat  (2.0 Mb)

Benitez N., Ford H., Bouwens R., Menanteau F., Blakeslee J., Gronwall C., Illingworth G., Meurer G., Broadhurst T.J., Clampin M., Franx M., Hartig G.F., Magee D., Sirianni M., Ardila D.R., Bartko F., Brown R.A., Burrows C.J., Cheng E.S., Cross N.J.G., Feldman P.D., Golimowski D.A., Infante L., Kimble R.A., Krist J.E., Lesser M.P., Levay Z., Martel A.R., Miley G.K., Postman M., Rosati P., Sparks W.B., Tran H.D., Tsvetanov Z.I., White R.L., Zheng W. Astrophys. J. Suppl. Ser., 150, 1 (2004) Faint galaxies in Deep Advanced Camera for Surveys observations.

J/ApJS/150/19/  table3.dat table6.dat table7.dat table8.dat table9.dat  (292Kb)

Kim D.-W., Cameron R.A., Drake J.J., Evans N.R., Freeman P., Gaetz T.J., Ghosh H., Green P.J., Harnden F.R.Jr, Karovska M., Kashyap V., Maksym P.W., Ratzlaff P.W., Schlegel E.M., Silverman J.D., Tananbaum H.D., Vikhlinin A.A., Wilkes B.J., Grimes J.P. Astrophys. J. Suppl. Ser., 150, 19 (2004) Chandra Multiwavelength Project. I. First X-ray source catalog.

J/ApJS/150/165/  table1.dat table2.dat table3.dat table4a.dat table4b.dat  (1.3 Mb)

Kuraszkiewicz J.K., Green P.J., Crenshaw D.M., Dunn J., Forster K., Vestergaard M., Aldcroft T.L. Astrophys. J. Suppl. Ser., 150, 165 (2004) Emission line properties of active galactic nuclei from a Post-COSTAR Hubble Space Telescope Faint Object Spectrograph spectral atlas.

J/ApJS/150/181/  table2.dat  (13Kb)

Eracleous M., Halpern J.P. Astrophys. J. Suppl. Ser., 150, 181 (2004) Accurate redshifts and classifications for 110 radio-loud active galactic nuclei

J/ApJS/150/343/  table2.dat  (44Kb)

Engelke C.W., Kraemer K.E., Price S.D. Astrophys. J. Suppl. Ser., 150, 343 (2004) A uniform database of 2.2-16.5μm spectra from the ISOCAM CVF spectrometer.

J/ApJS/150/367/  table5.dat table6.dat table7.dat table8.dat table9.dat  (71Kb)

Peng E.W., Ford H.C., Freeman K.C. Astrophys. J. Suppl. Ser., 150, 367-385 (2004) The globular cluster system of NGC 5128. I. Survey and catalogs.

J/ApJS/150/417/  table2.dat  (51Kb)

Cohen A.S., Rottgering H.J.A., Jarvis M.J., Kassim N.E., Lazio T.J.W. Astrophys. J. Suppl. Ser., 150, 417-430 (2004) A deep, high-resolution survey at 74 MHz.

J/ApJS/150/455/  table1.dat table2.dat table3.dat table4.dat  (89Kb)

Gould A., Chaname J. Astrophys. J. Suppl. Ser., 150, 455-464 (2004) New HIPPARCOS-based parallaxes for 424 faint stars.

J/ApJS/151/103 See V/116/

J/ApJS/151/271/  table1.dat  (18Kb)

Heiles C., Troland T.H. Astrophys. J. Suppl. Ser., 151, 271-297 (2004) The Millennium Arecibo 21 centimeter Absorption-Line Survey. III. Techniques for spectral polarization and results for Stokes V.

J/ApJS/151/299/  table5.dat  (151Kb)

Hodge T.M., Kraemer K.E., Price S.D., Walker H.J. Astrophys. J. Suppl. Ser., 151, 299-312 (2004) Classification of spectra from the Infrared Space Observatory PHT-S database.

J/ApJS/151/313/  table5.dat  (34Kb)

Pan K., Federman S.R., Cunha K., Smith V.V., Welty D.E. Astrophys. J. Suppl. Ser., 151, 313-343 (2004) Cloud structure and physical conditions in star-forming regions from optical observations. I. Data and component structure.

J/ApJS/151/387/  position.dat refs.dat table2.dat table3.dat table4.dat table5.dat  (110Kb)

Ivanov V.D., Rieke M.J., Engelbracht C.W., Alonso-herrero A., Rieke G.H., Luhman K.L. Astrophys. J. Suppl. Ser., 151, 387-397 (2004) A medium-resolution near-infrared spectral library of late-type stars. I.

J/ApJS/152/103/  table1.dat  (62.9 Mb)

Gould A., Kollmeier J.A. Astrophys. J. Suppl. Ser., 152, 103-111 (2004) Proper-motion catalog from SDSS intersection with USNO-B.

J/ApJS/152/201/  classes.dat table3.dat  (14Kb)

Gupta R., Singh H.P., Volk K., Kwok S. Astrophys. J. Suppl. Ser., 152, 201-209 (2004) Automated classification of 2000 bright IRAS sources.

J/ApJS/152/251/  table2.dat table3.dat table4.dat  (357Kb)

Valdes F., Gupta R., Rose J.A., Singh H.P., Bell D.J. Astrophys. J. Suppl. Ser., 152, 251-259 (2004) The Indo-US library of coude feed stellar spectra.

J/ApJS/152/261/  table1.dat table2.dat  (827Kb)

Wright J.T., Marcy G.W., Butler R.P., Vogt S.S. Astrophys. J. Suppl. Ser., 152, 261-295 (2004) Chromospheric Ca II emission in nearby F, G, K, and M stars.

J/ApJS/153/119/  notes.dat refs.dat table2.dat tablea1.dat tableb1.dat tablec1.dat  (69Kb)

Mitchell K.J., Usher P.D. Astrophys. J. Suppl. Ser., 153, 119-157 (2004) Faint blue objects at high galactic latitude. VIII. Performance characteristics of the US survey.

J/ApJS/153/223/  table1.dat  (11Kb)

Cote P., Blakeslee J.P., Ferrarese L., Jordan A., Mei S., Merritt D., Milosavljevic M., Peng E.W., Tonry J.L., West M.J. Astrophys. J. Suppl. Ser., 153, 223-242 (2004) The ACS Virgo Cluster Survey. I. Introduction to the survey.

J/ApJS/153/397/  table2.dat  (919Kb)

Barrientos L.F., Schade D., Lopez-Cruz O., Quintana H. Astrophys. J. Suppl. Ser., 153, 397-409 (2004) Photometric and morphological properties of nearby cluster galaxies.

J/ApJS/153/429/  table1.dat table2.dat table3.dat table4.dat table6.dat  (305Kb)

Kniazev A.Y., Pustilnik S.A., Grebel E.K., Lee H., Pramskij A.G. Astrophys. J. Suppl. Ser., 153, 429-445 (2004)

Strong emission line HII galaxies in the Sloan Digital Sky Survey. I. Catalog of DR1 objects with oxygen abundances from T_{e} measurements.

J/ApJS/154/155/  table1.dat  (9Kb)

Alonso-Herrero A., Perez-Gonzalez P.G., Rigby J., Rieke G.H., Le Floc’h E., Barmby P., Page M.J., Papovich C., Dole H., Egami E., Huang J.-S., Rigopoulou D., Cristobal-Hornillos D., Eliche-Moral C., Balcells M., Prieto M., Erwin P., Engelbracht C.W., Gordon K.D., Werner M., Willner S.P., Fazio G.G., Frayer D., Hines D., Kelly D., Latter W., Misselt K., Miyazaki S., Morrison J., Rieke M.J., Wilson G. Astrophys. J. Suppl. Ser., 154, 155-159 (2004) The nature of luminous X-ray sources with mid-infrared counterparts.

J/ApJS/154/509/  table1.dat  (33Kb)

Jordan A., Blakeslee J.P., Peng E.W., Mei S., Cote P., Ferrarese L., Tonry J.L., Merritt D., Milosavljevic M., West M.J. Astrophys. J. Suppl. Ser., 154, 509-517 (2004) The ACS Virgo Cluster Survey. II. Data reduction procedures.

J/ApJS/154/519/  table2.dat  (24Kb)

Swartz D.A., Ghosh K.K., Tennant A.F., Wu K. Astrophys. J. Suppl. Ser., 154, 519-539 (2004) The ultraluminous X-ray source population from the Chandra archive of galaxies.

J/ApJS/154/553/  table1.dat table2.dat table3.dat table4.dat table5.dat table7.dat  (66Kb)

Sewilo M., Watson C., Araya E., Churchwell E., Hofner P., Kurtz S. Astrophys. J. Suppl. Ser., 154, 553-578 (2004) Resolution of distance ambiguities of inner galaxy massive star formation regions. II.

J/ApJS/154/585/  table4.dat  (35Kb)

Harmon B.A., Wilson C.A., Fishman G.J., Connaughton V., Henze W., Paciesas W.S., Finger M.H., McCollough M.L., Sahi M., Peterson B., Shrader C.R., Grindlay J.E., Barret D. Astrophys. J. Suppl. Ser., 154, 585-622 (2004) The burst and transient source experiment (BATSE) earth occultation catalog of low-energy gamma-ray sources.

J/ApJS/154/623/  table2.dat  (119Kb)

Mould J., Saha A., Hughes S. Astrophys. J. Suppl. Ser., 154, 623-631 (2004) The period-luminosity relation for long-period variables in M31.

J/ApJS/154/673/  table1.dat table4.dat table5.dat table8.dat  (18.1 Mb)

Smith B.J., Price S.D., Baker R.I. Astrophys. J. Suppl. Ser., 154, 673 (2004) The COBE DIRBE Point Source Catalog.

J/ApJS/155/1/  table6.dat  (94Kb)

Sullivan M., Hopkins A.M., Afonso J., Georgakakis A., Chan B., Cram L.E., Mobasher B., Almeida C. Astrophys. J. Suppl. Ser., 155, 1-13 (2004) The Phoenix Deep Survey: optical and near-infrared imaging catalogs.

J/ApJS/155/33/  table1.dat table2.dat  (27Kb)

Scott W.K., Fomalont E.B., Horiuchi S., Lovell J.E.J., Moellenbrock G.A., Dodson R.G., Edwards P.G., Coldwell G.V., Fodor S., Frey S., Gurvits L.I., Hirabayashi H., Lister M.L., Mosoni L., Murata Y., Paragi Z., Piner B.G., Shen Z.-Q., Taylor A.R., Tingay S.J. Astrophys. J. Suppl. Ser., 155, 33-72 (2004) The VSOP 5 GHz active galactic nucleus survey. III. Imaging results for the first 102 sources.

J/ApJS/155/73/  table1.dat  (26Kb)

Zheng W., Mikles V.J., Mainieri V., Hasinger G., Rosati P., Wolf C., Norman C., Szokoly G., Gilli R., Tozzi P., Wang J.X., Zirm A., Giacconi R. Astrophys. J. Suppl. Ser., 155, 73-87 (2004) Photometric redshift of X-ray sources in the Chandra Deep Field-South.

J/ApJS/155/89/  table3.dat  (62Kb)

Gelfand J.D., Lazio T.J.W., Gaensler B.M. Astrophys. J. Suppl. Ser., 155, 89-100 (2004) A wide-field, low-frequency radio survey of the field of M31. I. Construction and statistical analysis of the source catalog.

J/ApJS/155/123/  refs.dat table1.dat table3.dat  (18Kb)

Alvarez C., Feldt M., Henning T., Puga E., Brandner W., Stecklum B. Astrophys. J. Suppl. Ser., 155, 123-148 (2004)

Near-infrared subarcsecond observations of ultracompact H II regions.

J/ApJS/155/149/  refs.dat table5.dat table6.dat table7.dat  (27Kb)

Kurtz S., Hofner P., Alvarez C.V. Astrophys. J. Suppl. Ser., 155, 149-165 (2004) A catalog of CH_{3}OH 7_{0}-6_{1} A^{+} maser sources in massive star-forming regions.

J/ApJS/155/175/  table1.dat  (58Kb)

Abt H.A. Astrophys. J. Suppl. Ser., 155, 175-177 (2004) Spectral classification of stars in a Supplement to the bright star catalogue.

J/ApJS/155/257/  table1.dat  (27.2 Mb)

Richards G.T., Nichol R.C., Gray A.G., Brunner R.J., Lupton R.H., Vanden Berk D.E., Chong S.S., Weinstein M.A., Schneider D.P., Anderson S.F., Munn J.A., Harris H.C., Strauss M.A., Fan X., Gunn J.E., Ivezic Z., York D.G., Brinkmann J., Moore A.W. Astrophys. J. Suppl. Ser., 155, 257-269 (2004) Efficient photometric selection of quasars from the Sloan Digital Sky Survey: 100000 z 3 quasars from data release one.

J/ApJS/155/271/  table4.dat table5.dat table8.dat table9.dat  (81Kb)

Szokoly G.P., Bergeron J., Hasinger G., Lehmann I., Kewley L., Mainieri V., Nonino M., Rosati P., Giacconi R., Gilli R., Gilmozzi R., Norman C., Romaniello M., Schreier E., Tozzi P., Wang J.X., Zheng W., Zirm A. Astrophys. J. Suppl. Ser., 155, 271-349 (2004) The Chandra Deep Field-South: Optical spectroscopy. I..

J/ApJS/155/351/  table3.dat  (27Kb)

Sembach K.R., Tripp T.M., Savage B.D., Richter P. Astrophys. J. Suppl. Ser., 155, 351-393 (2004) Physical properties and baryonic content of low-redshift intergalactic Lyαand O VI absorption line systems: the PG 1116+215 sight line.

J/ApJS/155/401/  notes.dat table1.dat table5.dat  (52Kb)

Tanabe T., Kucinskas A., Nakada Y., Onaka T., Sauvage M. Astrophys. J. Suppl. Ser., 155, 401-419 (2004) ISOCAM observations of globular clusters in the Magellanic clouds: the data.

J/ApJS/155/421/  refs.dat table1.dat table2.dat table3.dat table4.dat  (63Kb)

Yusef-Zadeh F., Hewitt J.W., Cotton W. Astrophys. J. Suppl. Ser., 155, 421-550 (2004) A 20 centimeter survey of the Galactic Center region. I. Detection of numerous linear filaments.

J/ApJS/155/577/  tables.dat  (35Kb)

Mcgrath E.J., Goss W.M., De Pree C.G. Astrophys. J. Suppl. Ser., 155, 577-593 (2004) H_{2}O masers in W49 north and Sagittarius B2.

J/ApJS/155/595/  table1.dat table3.dat  (38Kb)

Deacon R.M., Chapman J.M., Green A.J. Astrophys. J. Suppl. Ser., 155, 595-622 (2004) OH maser observations of likely planetary nebulae precursors.

J/ApJS/155/651/  table1.dat table5.dat table6.dat  (168Kb)

Herwig F. Astrophys. J. Suppl. Ser., 155, 651-666 (2004) Evolution and yields of extremely metal-poor intermediate-mass stars.

J/ApJS/156/35/  table3.dat  (124Kb)

Eckart M.E., Laird E.S., Stern D., Mao P.H., Helfand D.J., Harrison F.A. Astrophys. J. Suppl. Ser., 156, 35-45 (2005) The serendipitous extragalactic X-ray source identification (SEXSI) program. II. Optical imaging.

J/ApJS/156/47/  table1.dat  (38Kb)

Liebert J., Bergeron P., Holberg J.B. Astrophys. J. Suppl. Ser., 156, 47-68 (2005) The formation rate and mass and luminosity functions of DA white dwarfs from the Palomar Green survey.

J/ApJS/156/69/  table5.dat table6.dat  (38Kb)

Weidenspointner G., Harris M.J., Sturner S., Teegarden B.J., Ferguson C. Astrophys. J. Suppl. Ser., 156, 69-91 (2005) MGGPOD: a Monte Carlo suite for modeling instrumental line and continuum backgrounds in gamma-ray

astronomy.

J/ApJS/156/105/  table3.dat table4.dat  (96Kb)

Gu M.F. Astrophys. J. Suppl. Ser., 156, 105-110 (2005) Wavelengths of 2l– 3l’ transitions in L-shell ions of iron and nickel: a combined configuration interaction and many-body perturbation theory approach.

J/ApJS/156/127/ notes.dat table3.dat table4.dat table5.dat table6.dat table7.dat table8.dat table9.dat table10.dat table11.dat table12.dat table13.dat table14.dat table15.dat table16.dat table17.dat (105Kb) Comito C., Schilke P., Phillips T.G., Lis D.C., Motte F., Mehringer D. Astrophys. J. Suppl. Ser., 156, 127-167 (2005) A molecular line survey of Orion KL in the 350 micron band.

J/ApJS/156/169/  refs.dat table1.dat table2.dat table3.dat table4.dat  (50Kb)

Froebrich D. Astrophys. J. Suppl. Ser., 156, 169-177 (2005) Which are the youngest protostars? Determining properties of confirmed and candidate class 0 sources by broadband photometry.

J/ApJS/156/217/  notes.dat table1.dat table2.dat table3.dat  (38Kb)

Hurley K., Stern B., Kommers J., Cline T., Mazets E., Golenetskii S., Trombka J., McClanahan T., Goldsten J., Feroci M., Frontera F., Guidorzi C., Montanari E., Lewin W., Meegan C., Fishman G., Kouveliotou C., Sinha S., Seetha S. Astrophys. J. Suppl. Ser., 156, 217-226 (2005) The interplanetary network supplement to the BATSE catalogs of untriggered cosmic gamma-ray bursts.

J/ApJS/156/227/  table3.dat  (31Kb)

Barnes III T.G., Jeffery E.J., Montemayor T.J., Skillen I. Astrophys. J. Suppl. Ser., 156, 227-235 (2005) Radial velocities of Galactic Cepheids.

J/ApJS/156/345/  table1.dat table2.dat table3.dat  (46Kb)

Gil De Paz A., Madore B.F. Astrophys. J. Suppl. Ser., 156, 345-360 (2005) Palomar/Las Campanas imaging atlas of blue compact dwarf galaxies. II. Surface photometry and the properties of the underlying stellar population.

J/ApJS/157/1/  table5.dat table6.dat  (47Kb)

Gladders M.D., Yee H.K.C. Astrophys. J. Suppl. Ser., 157, 1-29 (2005) The Red-Sequence Cluster Survey. I. The survey and cluster catalogs for patches RCS 0926+37 and RCS 1327+29.

J/ApJS/157/59/  table1.dat table2.dat table3.dat  (151Kb)

Liu J.-F., Bregman J.N. Astrophys. J. Suppl. Ser., 157, 59-125 (2005) Ultraluminous X-ray sources in nearby galaxies from ROSAT High Resolution Imager observations. I. Data analysis.

J/ApJS/157/138/  refs.dat table2.dat  (132Kb)

Nielsen K.E., Gull T.R., Kober G.V. Astrophys. J. Suppl. Ser., 157, 138-146 (2005) The ultraviolet spectrum of eta Carinae: investigation of the ejecta absorption.

J/ApJS/157/175/  table1.dat table2.dat table3.dat  (56Kb)

Koo D.C., Simard L., Willmer C.N.A., Gebhardt K., Bouwens R.J., Kauffmann G., Crosby T., Faber S.M., Harker J., Sarajedini V.L., Vogt N.P., Weiner B.J., Phillips A.J., Im M., Wu K.L. Astrophys. J. Suppl. Ser., 157, 175-217 (2005) The DEEP Groth Strip Survey. VIII. The evolution of luminous field bulges at redshift z 1.

J/ApJS/157/228/  table2.dat table3.dat table4.dat table5.dat  (227Kb)

Saintonge A., Schade D., Ellingson E., Yee H.K.C., Carlberg R.G. Astrophys. J. Suppl. Ser., 157, 228-250 (2005) Catalog of galaxy morphology in four rich clusters: luminosity evolution of disk galaxies at 0.33 z 0.83.

J/ApJS/157/324/  table4.dat table5 9.dat  (23Kb)

Thompson D.J., Bertsch D.L., O’Neal R.H.JR Astrophys. J. Suppl. Ser., 157, 324-334 (2005) The highest-energy photons seen by the Energetic Gamma Ray Experiment Telescope (EGRET) on the Compton Gamma Ray Observatory.

J/ApJS/157/371/  table2.dat  (52Kb)

Robertson-tessi M., Garnett D.R. Astrophys. J. Suppl. Ser., 157, 371-395 (2005) Recombination line versus forbidden line abundances in planetary nebulae.

J/ApJS/158/139/  table1.dat table2.dat table3.dat table4.dat table5.dat table6.dat  (36Kb)

Keel W.C., Irby B.K., May A., Miley G.K., Golombek D., De Grijp M.H.K., Gallimore J.F. Astrophys. J. Suppl. Ser., 158, 139-160 (2005) An atlas of warm active galactic nuclei and starbursts from the IRAS deep fields.

J/ApJS/158/161/  table1.dat table3.dat  (193.1 Mb)

Hsieh B.C., Yee H.K.C., Lin H., Gladders M.D. Astrophys. J. Suppl. Ser., 158, 161-177 (2005) A photometric redshift galaxy catalog from the Red-Sequence Cluster Survey.

J/ApJS/158/188/  table1.dat  (265Kb)

Widicus Weaver S.L., Butler R.A.H., Drouin B.J., Petkie D.T., Dyl K.A., De Lucia F.C., Blake G.A. Astrophys. J. Suppl. Ser., 158, 188-192 (2005) Millimeter-wave and vibrational state assignments for the rotational spectrum of glycolaldehyde.

J/ApJS/158/193/ refs1.dat refs2.dat refs3.dat refs4.dat table1.dat table2.dat table3.dat table4.dat (76Kb) Stankov A., Handler G. Astrophys. J. Suppl. Ser., 158, 193-216 (2005) Catalog of Galactic βCephei stars.

J/ApJS/159/41/  table1.dat  (1.3 Mb)

Vogt N.P., Koo D.C., Phillips A.C., Wu K., Faber S.M., Willmer C.N.A., Simard L., Weiner B.J., Illingworth G.D., Gebhardt K., Gronwall C., Guzman R., Im M., Sarajedini V., Groth E.J., Rhodes J., Brunner R., Connolly A., Szalay A., Kron R., Blandford R. Astrophys. J. Suppl. Ser., 159, 41-59 (2005) The DEEP Groth Strip Survey. I. The sample.

J/ApJS/159/100/  table6.dat table7.dat table8.dat  (15Kb)

Taylor B.J., Joner M.D. Astrophys. J. Suppl. Ser., 159, 100 (2005) A catalog of temperatures and red Cousins photometry for the Hyades.

J/ApJS/159/141/  table8.dat table9.dat  (242Kb)

Valenti J.A., Fischer D.A. Astrophys. J. Suppl. Ser., 159, 141-166 (2005) Spectroscopic properties of cool stars (SPOCS). I. 1040 F, G, and K dwarfs from Keck, Lick, and AAT planet search programs.

J/ApJS/159/242/  table1.dat table2.dat table3.dat table5.dat table6.dat table8.dat table9.dat  (27Kb)

Gelfand J.D., Lazio T.J.W., Gaensler B.M. Astrophys. J. Suppl. Ser., 159, 242-276 (2005) A wide-field, low-frequency radio image of the field of M31. II. Source classification and discussion

J/ApJS/159/282/  table1.dat  (32Kb)

Luo S.G., Condon J.J., Yin Q.F. Astrophys. J. Suppl. Ser., 159, 282-287 (2005) Radio identifications of recently discovered planetary nebulae.

J/ApJS/160/149 See VIII/77/

J/ApJS/160/220/ refs.dat table1.dat table2.dat table3.dat table4.dat table5.dat table6.dat table7.dat table8.dat table9.dat table10.dat table11.dat table12.dat table13.dat table14.dat table15.dat table16.dat table17.dat table18.dat table19.dat table20.dat table21.dat (179Kb) Fish V.L., Reid M.J., Argon A.L., Zheng X.-W. Astrophys. J. Suppl. Ser., 160, 220-271 (2005) Full-polarization observations of OH masers in massive star-forming regions. I. Data.

J/ApJS/160/272/  table1.dat  (13Kb)

Huggins P.J., Bachiller R., Planesas P., Forveille T., Cox P. Astrophys. J. Suppl. Ser., 160, 272-285 (2005) A CO survey of young planetary nebulae.

J/ApJS/160/319/ notes.dat table2.dat table4.dat table6.dat table7.dat table8.dat table9.dat table10.dat table11.dat table12.dat (1.2 Mb) Getman K.V., Flaccomio E., Broos P.S., Grosso N., Tsujimoto M., Townsley L., Garmire G.P., Kastner J., Li J., Harnden F.R.JR, Wolk S., Murray S.S., Lada C.J., Muench A.A., McCaughrean M.J., Meeus G., Damiani F., Micela G., Sciortino S., Bally J., Hillenbrand L.A., Herbst W., Preibisch T., Feigelson E.D. Astrophys. J. Suppl. Ser., 160, 319-352 (2005) Chandra Orion Ultradeep Project: observations and sources lists.

J/ApJS/160/353/  table1.dat table2.dat table3.dat table4.dat  (44Kb)

Getman K.V., Feigelson E.D., Grosso N., McCaughrean M.J., Micela G., Broos P., Garmire G., Townsley L. Astrophys. J. Suppl. Ser., 160, 353-378 (2005) Membership of the Orion nebula population from the Chandra Orion Ultradeep Project.

J/ApJS/160/450/  table1.dat  (24Kb)

Flaccomio E., Micela G., Sciortino S., Feigelson E.D., Herbst W., Favata F., Harnden F.R.JR, Vrtilek S.D. Astrophys. J. Suppl. Ser., 160, 450-468 (2005) Rotational modulation of X-ray emission in Orion Nebula young stars.

J/ApJS/160/511/  table1.dat table2.dat table3.dat  (32Kb)

Kastner J.H., Franz G., Grosso N., Bally J., McCaughrean M.J., Getman K., Feigelson E.D., Schulz N.S. Astrophys. J. Suppl. Ser., 160, 511-529 (2005) X-ray emission from Orion Nebula Cluster stars with circumstellar disks and jets.

J/ApJS/161/9/  table1.dat table3.dat  (456Kb)

Kenter A., Murray S.S., Forman W.R., Jones C., Green P., Kochanek C.S., Vikhlinin A., Fabricant D., Fazio G., Brand K., Brown M.J.I., Dey A., Jannuzi B.T., Najita J., Mcnamara B., Shields J., Rieke M. Astrophys. J. Suppl. Ser., 161, 9-20 (2005) XBootes: an X-ray survey of the NDWFS Bootes field. II. The X-ray source catalog.

J/ApJS/161/21/  table2.dat table3.dat table6.dat  (274Kb)

Lehmer B.D., Brandt W.N., Alexander D.M., Bauer F.E., Schneider D.P., Tozzi P., Bergeron J., Garmire G.P., Giacconi R., Gilli R., Hasinger G., Hornschemeier A.E., Koekemoer A.M., Mainieri V., Miyaji T., Nonino M., Rosati P., Silverman J.D., Szokoly G., Vignali C. Astrophys. J. Suppl. Ser., 161, 21-40 (2005) The Extended Chandra Deep Field-South survey: Chandra point-source catalogs.

J/ApJS/161/118/  refs.dat table1.dat table2.dat table4.dat  (1.5 Mb)

McSwain M.V., Gies D.R. Astrophys. J. Suppl. Ser., 161, 118-146 (2005) The evolutionary status of Be stars: results from a photometric study of southern open clusters.

J/ApJS/161/147/  table3.dat  (52Kb)

Thom C., Gibson B.K., Christlieb N. Astrophys. J. Suppl. Ser., 161, 147-153 (2005) H I spectra and associated moment maps for all lines of sight.

J/ApJS/161/154/  atlas/* table1.dat table3.dat  (1.1 Mb)

Hanson M.M., Kudritzki R.-P., Kenworthy M.A., Puls J., Tokunaga A.T. Astrophys. J. Suppl. Ser., 161, 154-170 (2005) A medium resolution near-infrared spectral atlas of O and early-B stars.

J/ApJS/161/185/  table1.dat table2.dat table4.dat  (208Kb)

Ueda Y., Ishisaki Y., Takahashi T., Makishima K., Ohashi T. Astrophys. J. Suppl. Ser., 161, 185-223 (2005) The ASCA medium sensitivity survey (The GIS catalog project): source catalog II.

J/ApJS/161/240/  table1.dat table3.dat table4.dat table5.dat table6.dat  (119Kb)

Thuan T.X., Izotov Y.I. Astrophys. J. Suppl. Ser., 161, 240-270 (2005) High-ionization emission in metal-deficient blue compact dwarf galaxies.

J/ApJS/161/271/  table3.dat table4.dat table5.dat table6.dat  (152Kb)

Grimm H.-J., Mcdowell J., Zezas A., Kim D.-W., Fabbiano G. Astrophys. J. Suppl. Ser., 161, 271-303 (2005) The X-ray binary population in M33. I. Source list and luminosity function.

J/ApJS/161/361/  table2.dat table4.dat table56.dat  (25Kb)

Klein R., Posselt B., Schreyer K., Forbrich J., Henning T. Astrophys. J. Suppl. Ser., 161, 361-393 (2005) A millimeter continuum survey for massive protoclusters in the outer Galaxy.

J/ApJS/161/394/  refs.dat table1.dat table3.dat table5.dat table6.dat  (59Kb)

Farihi J., Becklin E.E., Zuckerman B. Astrophys. J. Suppl. Ser., 161, 394-428 (2005) Low-luminosity companions to white dwarfs.

J/ApJS/161/444/  refs.dat table8.dat table9.dat  (146Kb)

Taylor B.J. Astrophys. J. Suppl. Ser., 161, 444-455 (2005) Statistical cataloging of archival data for luminosity class IV-V stars. III. The epoch 2004 [Fe/H] and temperature catalogs.

J/ApJS/162/1/  table7.dat  (25.7 Mb)

Gawiser E., Van Dokkum P.G., Herrera D., Maza J., Castander F.J., Infante L., Lira P., Quadri R., Toner R., Treister E., Urry C.M., Altmann M., Assef R., Christlein D., Coppi P.S., Duran M.F., Franx M., Galaz G., Huerta L., Liu C., Lopez S., Mendez R., Moore D.C., Rubio M., Ruiz M.T., Toft S., Yi S.K. Astrophys. J. Suppl. Ser., 162, 1-19 (2006) The Multiwavelength Survey by Yale-Chile (MUSYC): survey design and deep public UBVRIz’ images and catalogs of the Extended Hubble Deep Field-South.

J/ApJS/162/38 See II/267/

J/ApJS/162/49/  refs.dat table1.dat table2.dat table3.dat table4.dat table5.dat table6.dat  (140Kb)

Hunter D.A., Elmegreen B.G. Astrophys. J. Suppl. Ser., 162, 49-79 (2006) Broadband imaging of a large sample of irregular galaxies.

J/ApJS/162/80/  table1.dat table2.dat table3.dat table4.dat  (30Kb)

Kassin S.A., De Jong R.S., Pogge R.W. Astrophys. J. Suppl. Ser., 162, 80-96 (2006) Dark and baryonic matter in bright spiral galaxies. I. Near-infrared and optical broadband surface photometry of 30 galaxies.

J/ApJS/162/207/  stars.dat table2.dat  (105Kb)

Abt H.A., Willmarth D. Astrophys. J. Suppl. Ser., 162, 207-226 (2006) The secondaries of solar-type primaries. I. The radial velocities.

J/ApJS/162/227/  table1.dat table2.dat  (70Kb)

Lawler J.E., Den Hartog E.A., Sneden C., Cowan J.J. Astrophys. J. Suppl. Ser., 162, 227-260 (2006) Improved laboratory transition probabilities for Sm II and application to the samarium abundances of the Sun and three r-process-rich, metal-poor stars.

J/ApJS/162/304/  table4.dat  (73Kb)

Henry J.P., Mullis C.R., Voges W., Bohringer H., Briel U.G., Gioia I.M., Huchra J.P. Astrophys. J. Suppl. Ser., 162, 304-328 (2006) The ROSAT North Ecliptic Pole survey: the X-ray catalog.

J/ApJS/162/346/  refs.dat table4.dat table6.dat  (75Kb)

Rudolph A.L., Fich M., Bell G.R., Norsen T., Simpson J.P., Haas M.R., Erickson E.F. Astrophys. J. Suppl. Ser., 162, 346-374 (2006) Abundance gradients in the Galaxy.

J/ApJS/162/428/  table1.dat table3.dat  (723Kb)

Maeda A., De Lucia F.C., Herbst E., Pearson J.C., Riccobono J., Trosell E., Bohn R.K. Astrophys. J. Suppl. Ser., 162, 428-435 (2006) The millimeter- and submillimeter-wave spectrum of the Gt conformer of n-propanol (n-CH_{3}CH_{2}CH_{2}OH).

J/ApJS/163/80/  table3.dat table4.dat  (223Kb)

McDonald P., Seljak U., Burles S., Schlegel D.J., Weinberg D.H., Cen R., Shih D., Schaye J., Schneider D.P., Bahcall N.A., Briggs J.W., Brinkmann J., Brunner R.J., Fukugita M., Gunn J.E., Ivezic Z., Kent S., Lupton R.H., Vanden Berk D.E. Astrophys. J. Suppl. Ser., 163, 80-109 (2006) The Lyαforest power spectrum from the Sloan Digital Sky Survey.

J/ApJS/163/207/  table1.dat table2.dat table3.dat  (71Kb)

Tayal S.S. Astrophys. J. Suppl. Ser., 163, 207-223 (2006) New accurate oscillator strengths and electron excitation collision strengths for N I.

J/ApJS/163/306/  table1.dat table2.dat table3.dat  (126Kb)

Getman K.V., Feigelson E.D., Townsley L., Broos P., Garmire G., Tsujimoto M. Astrophys. J. Suppl. Ser., 163, 306-334 (2006) Chandra study of the Cepheus B star-forming region: stellar populations and the initial mass function.

J/ApJS/165/1/  table4.dat  (773Kb)

Trump J.R., Hall P.B., Reichard T.A., Richards G.T., Schneider D.P., Vanden Berk D.E., Knapp G.R., Anderson S.F., Fan X., Brinkman J., Kleinman S.J., Nitta A. Astrophys. J. Suppl. Ser., 165, 1-18 (2006) A catalog of broad absorption line quasars from the Sloan Digital Sky Survey third data release.

J/ApJS/165/19/  notes.dat table2.dat  (73Kb)

Eckart M.E., Stern D., Helfand D.J., Harrison F.A., Mao P.H., Yost S.A. Astrophys. J. Suppl. Ser., 165, 19-56 (2006) The Serendipitous Extragalactic X-ray Source Identification (SEXSI) program. III. Optical spectroscopy.

J/ApJS/166/526/  table1.dat table3.dat table4.dat  (28Kb)

Xu Y., Reid M.J., Menten K.M., Zheng X.W. Astrophys. J. Suppl. Ser., 166, 526-533 (2006) Search for compact extragalactic radio sources near massive star-forming regions.

J/ApJS/166/549/  table2.dat  (5.5 Mb)

Gouliermis D.A., Dolphin A.E., Brandner W., Henning T. Astrophys. J. Suppl. 166, 549-556 (2006) The star-forming region NGC 346 in the small Magellanic cloud with Hubble Space telescope ACS observations. I. Photometry.

J/ApJS/167/315/  table1.dat table3.dat table4.dat table5.dat  (41Kb)

Nahar S.N. Astrophys. J. Suppl. Ser., 167, 315-333 (2006) Electron-ion recombination rate coefficients and photoionization cross sections for astrophysically abundant elements. XII. Na IX, Na X, Mg X, and Mg XI for ultraviolet and X-ray modeling.

J/ApJS/169/430/  table2.dat table3.dat  (257Kb)

Robinson S.E., Ammons S.M., Kretke K.A., Strader J., Wertheimer J.G., Fischer D.A., Laughlin G. Astrophys. J. Suppl. Ser., 169, 430-438 (2007) The N2K consortium. VII. Atmospheric parameters of 1907 metal-rich stars: finding planet-search targets.

J/ApJS/170/95/  table4.dat  (1.0 Mb)

Chiboucas K., Mateo M. Astrophys. J. Suppl. Ser., 170, 95-107 (2007) The luminosity function of nearby galaxy clusters. I. A catalog of galaxies within the central 0.83 square degrees of the Centaurus cluster.

J/ApJS/170/175/  notes.dat table1.dat types.dat  (54Kb)

Bird A.J., Malizia A., Bazzano A., Barlow E.J., Bassani L., Hill A.B., Belanger G., Capitanio F., Clark D.J., Dean A.J., Fiocchi M., Gotz D., Lebrun F., Molina M., Produit N., Renaud M., Sguera V., Stephen J.B., Terrier R., Ubertini P., Walter R., Winkler C., Zurita J. Astrophys. J. Suppl. Ser., 170, 175-186 (2007) The third IBIS/ISGRI soft gamma-ray survey catalog.

J/BaltA/5/1/  notes.dat tables.dat  (171Kb)

Bartkevicius A., Lazauskaite R. Baltic Astronomy, 5, 1 (1996) Classification of population II stars in the Vilnius photometric system. I. Methods

J/BaltA/7/369/  table2.dat table4.dat  (31Kb)

Boyle R.P., Kazlauskas A., Vansevicius V., Straizys V., Vrba F.J., Sudzius J., Smriglio F. Baltic Astronomy, 7, 369 (1998) CCD photometry of the M67 cluster in the Vilnius photometric system

J/BaltA/8/325/  notes.dat table1.dat table2.dat  (56Kb)

Bartkevicius A., Sperauskas J. Baltic Astronomy, 8, 325 (1999) Radial velocities of Population II Stars

J/BaltA/9/646/  catalog.dat  (137Kb)

Kharchenko N.V., Kilpio E.Yu., Malkov O.Yu., Schilbach E. Baltic Astronomy, 9, 646 (2000) The Compiled Catalogue of the astronomical data of the Mira-type stars.

J/BaltA/10/319/  notes.dat table2.dat table4.dat  (40Kb)

Straizys V., Cernis K., Bartasiute S. Baltic Astronomy, 10, 319-347 (2001) Photometric investigation of the California nebula region: catalogs

J/BaltA/10/349/  notes.dat table1.dat table4.dat  (42Kb)

Zdanavicius J., Cernis K., Zdanavicius K., Straizys V. Baltic Astronomy, 10, 349 (2001) Photometric classification of stars and the interstellar extinction near the Camelopardalis and Perseus border.

J/BaltA/10/481/ refs.dat table1.dat table2.dat table3.dat table4.dat table5.dat table6.dat table7.dat table8.dat table9.dat table10.dat table13a.dat table13b.dat (252Kb) Bartkevicius A., Gudas A. Baltic Astronomy, 10, 481 (2001) Kinematics of Hipparcos visual binaries. I. Stars with orbital solutions.

J/BaltA/10/589/  table1.dat  (147Kb)

Adelman S.J. Baltic Astronomy, 10, 589 (2001) Stars with the largest Hipparcos photometric amplitudes.

J/BaltA/11/1/  table1.dat table2.dat table3.dat  (98Kb)

Wegner W. Baltic Astronomy, 11, 1 (2002) Atlas of interstellar extinction curves of OB stars covering the whole available wavelength range.

J/BaltA/11/75/  notes.dat table1.dat  (45Kb)

Zdanavicius J., Zdanavicius K. Baltic Astronomy, 11, 75 (2002) Photometry and classification of stars along the Camelopardalis and Perseus border.

J/BaltA/11/153/ table1.dat table2.dat table3.dat table4.dat table5.dat table6.dat table7.dat table8.dat table9.dat table12.dat table13.dat (318Kb) Bartkevicius A., Gudas A. Baltic Astronomy, 11, 153 (2002) Kinematics of Hipparcos visual binaries. II. Stars with ground-based orbital solutions.

J/BaltA/11/219/  notes.dat table1.dat  (24Kb)

Kazlauskas A., Cernis K., Laugalys V., Straizys V. Baltic Astron., 11, 219 (2002) Photometric investigation of the MBM 12 molecular cloud area in Aries. I. Photoelectric photometry.

J/BaltA/11/231/  notes.dat table1.dat  (23Kb)

Straizys V., Cernis K., Kazlauskas A., Laugalys V. Baltic Astron., 11, 231 (2002) Photometric investigation of the MBM 12 molecular cloud area in Aries. II. Cloud distance.

J/BaltA/11/321/  notes.dat table1.dat  (56Kb)

Straizys V., Cernis K., Bartasiute S. Baltic Astron., 11, 321 (2002) Photometric investigation of the galaxy in the direction of Serpens cauda. Photoelectric photometry.

J/BaltA/11/417/  notes.dat table2.dat table3.dat table4.dat table5.dat  (49Kb)

Straizys V., Bartasiute S., Cernis K. Baltic Astron., 11, 417 (2002) Photometric investigation of the galaxy in the direction of Serpens cauda. A catalog of extinctions and distances.

J/BaltA/11/441/  table3.dat table4.dat table5.dat  (28Kb)

Zdanavicius J., Zdanavicius K. Baltic Astronomy, 11, 441 (2002) Interstellar extinction along the Camelopardalis and Perseus border.

J/BaltA/11/465/  table1.dat  (19Kb)

Sperauskas J., Aslan Z., Bartasiute S., Boyle R.P. Baltic Astronomy, 11, 465 (2002) Radial velocities of stars in the mega proper motion fields at high galactic latitudes.

J/BaltA/12/301/  notes.dat table1.dat table2.dat  (31Kb)

Cernis K., Straizys V. Baltic Astronomy, 12, 301-321 (2003) Interstellar extinction in the direction of the Barnard 1 dark cloud in Perseus.

J/BaltA/12/323/  table1.dat table2.dat table3.dat  (46Kb)

Straizys V., Kazlauskas A., Cerniauskas A., Boyle R.P., Vrba F.J., Philip A.G.D., Laugalys V., Cernis K., Smriglio F.

Baltic Astronomy, 12, 323-351 (2003) Overlapping open clusters NGC 1750 and NGC 1758 behind the Taurus dark clouds. II. CCD photometry in the Vilnius system.

J/BaltA/13/1/  table2.dat table4.dat  (46Kb)

Laugalys V., Kazlauskas A., Boyle R.P., Vrba F.J., Philip A.G.D., Straizys V. Baltic Astronomy, 13, 1-33 (2004) CCD photometry of the M 67 cluster in the Vilnius system. II. New photometry of high accuracy.

J/BaltA/13/555/  table1.dat table2.dat table3.dat table4.dat  (30Kb)

Zdanavicius J., Zdanavicius K., Straizys V., Kazlauskas A., Cernis K., Chen C.W., Chen W.P., Boyle R.P., Tautvaisiene G. Baltic Astronomy, 13, 555-572 (2004) Seven-color photometry of the open cluster NGC 2395 area.

J/BaltA/14/1/  notes.dat table2.dat  (137Kb)

Zdanavicius J., Zdanavicius K. Baltic Astron., 14, 1-40 (2005) CCD photometry and classification of stars in a Camelopardalis area.

J/BaltA/14/41/  table2.dat table4.dat table6.dat table8.dat tablea.dat  (57Kb)

Adelman S.J., Dennis J.W.III Baltic Astronomy, 14, 41 (2005) BVRI photometry of the S-type asymptotic branch giant stars RS Cnc, ST Her, OP Her, and HR Peg.

J/BaltA/14/51/  abund.dat tablea1.dat tablea1.tex  (62Kb)

Yuce K. Baltic Astronomy, 14, 51 (2005) Elemental abundances of 4 Lacertae and νCephei

J/MNRAS/120/287 See II/41/

J/MNRAS/121/103 See VII/216/

J/MNRAS/122/363/ table3.dat table4.dat (11Kb) Cuffey J. Mon. Not. R. Astron. Soc. 122, 363 (1961) NGC 7492

J/MNRAS/123/521 See II/40/

J/MNRAS/128/261 See II/63A/

J/MNRAS/128/435 See II/24/

J/MNRAS/128/475 See II/25/

J/MNRAS/131/467 See II/22/

J/MNRAS/140/21 See III/25/

J/MNRAS/154/343 See II/19/

J/MNRAS/169/171 See II/63A/

J/MNRAS/171/475/ 5c5.dat (26Kb) Pearson T.J. Mon. Not. R. Astron. Soc. 171, 475 (1975) The 5C5 survey of radio sources.

J/MNRAS/174/267 See VII/14/

J/MNRAS/177/499 See II/178/

J/MNRAS/179/635 See VII/8/

J/MNRAS/182/273/ 5c6.dat 5c7.dat (69Kb) Pearson T.J. Mon. Not. R. Astron. Soc. 182, 273 (1978) The 5C6 and 5C7 surveys of radio sources.

J/MNRAS/182/489 See VII/19/

J/MNRAS/185/613 See VII/23/

J/MNRAS/194/403 See II/73/

J/MNRAS/194/693 See VIII/2/ J/MNRAS/200/747/ 5c12.dat (31Kb)

Benn C.R., Grueff G., Vigotti M., Wall J.V. Mon. Not. R. Astron. Soc. 200, 747 (1982) A deep radio/optical survey near the North Galactic Pole. I. The 5C12 catalogue.

J/MNRAS/201/495/ strong.fit (5.8 Mb) Strong A.W., Riley P.A., Osborne J.L. Mon. Not. R. Astron. Soc. 201, 495 (1982) A Neutral Hydrogen Survey of the Southern Galactic Plane for —b— 10 deg.

J/MNRAS/201/1139 See II/91/

J/MNRAS/202/113 See VII/42A/

J/MNRAS/208/911 See I/85/

J/MNRAS/211/637/ table3 (16Kb) Murphy H.P. Mon. Not. R. Astron. Soc. 211, 637 (1984) CCD camera observations of nearby rich clusters - II. Techniques and results for A85

J/MNRAS/216/173/ 11cmasky.dat refs.dat (41Kb) Wall J.V., Peacock J.A. Mon. Not. R. Astron. Soc. 216, 173 (1985) Bright extragalactic radio sources at 2.7 GHz- III. The all-sky catalogue.

J/MNRAS/217/717 See VIII/18/

J/MNRAS/223/279 See VII/169/

J/MNRAS/223/629 See I/129/

J/MNRAS/224/61/  table3.dat  (19Kb)

Walker A.R., Laney C.D. Mon. Not. R. Astron. Soc. 224, 61 (1987) CCD photometry of galactic clusters containing Cepheid variables - IV. NGC 6649

J/MNRAS/224/453/ table3.dat (37Kb) Colless M., Hewett P. Mon. Not. R. Astron. Soc. 224, 453 (1987) The dynamics of rich clusters - I. Velocity data

J/MNRAS/227/563 See VII/169/

J/MNRAS/227/705 See VIII/45/

J/MNRAS/229/589/ table1.dat (160Kb) Purvis A., Tappin S.J., Rees W.G., Hewish A., Duffett-Smith P.J. Mon. Not. R. Astron. Soc. 229, 589 (1987) The Cambridge IPS Survey at 81.5 MHz

J/MNRAS/234/583/ table1a table1b table1c table1d (73Kb) Monk D.J., Barlow M.J., Clegg R.E.S. Mon. Not. R. Astron. Soc. 234, 583 (1988) An optical spectrophotometric survey of abundances in Magellanic Cloud planetary nebulae

J/MNRAS/234/919 See VIII/21/

J/MNRAS/235/1313/ table (69Kb) Slee O.B., Siegman B.C. Mon. Not. R. Astron. Soc. 235, 1313 (1988) A survey for metrewave variability extragalactic radio sources.

J/MNRAS/236/207/ table5.dat (77Kb) Metcalfe N., Fong R., Shanks T., Kilkenny D. Mon. Not. R. Astron. Soc. 236, 207 (1989) An extended galaxy redshift survey - I. The catalogue

J/MNRAS/237/495 See III/153/

J/MNRAS/238/709/ appena (13Kb) Stobie R.S., Ishida K., Peacock J.A. Mon. Not. R. Astron. Soc. 238, 709 (1989) Distance errors and the stellar luminosity function

J/MNRAS/238/1121 See II/160/ J/MNRAS/238/1171/ pks opir.dat refs.dat (36Kb)

Dunlop J.S., Peacock J.A., Savage A., Lilly S.J., Heasley J.N., Simon A.J.B. Mon. Not. R. Astron. Soc. 238, 1171 (1989) The Parkes selected regions : deep optical and infrared observations of radio galaxies and quasars at high redshifts.

J/MNRAS/239/305 See III/151/

J/MNRAS/239/459/ a2197.dat a2199.dat (458Kb) Dixon K.L., Godwin J.G., Peach J.V. Mon. Not. R. Astron. Soc., 239, 459-478 (1989) The clusters of galaxies A2197 and A2199.

J/MNRAS/239/605 See III/151/

J/MNRAS/240/785/ appendix.dat (56Kb) Millington S.J.C., Peach J.V. Mon. Not. R. Astron. Soc. 240, 785 (1989) Multicolour photometry of the cluster of galaxies Sersic 129-01

J/MNRAS/242/167 See I/185/

J/MNRAS/243/1/  table6.dat table7.dat table8.dat  (96Kb)

Boyle B.J., Fong R., Shanks T., Peterson B.A. Mon. Not. R. Astron. Soc. 243, 1 (1990) A catalogue of faint, UV-excess objects.

J/MNRAS/243/390/ table1 (33Kb) KARACHENTSEV I.D., KOPYLOV A.I. Mon. Not. R. Astron. Soc. 243, 390 (1990) Galaxy redshift survey in a narrow strip crossing the Coma cluster

J/MNRAS/244/408/ table3 (17Kb) COLLESS M., ELLIS R.S., TAYLOR K., HOOK R.N. Mon. Not. R. Astron. Soc. 244, 408 (1990) The LDSS deep redshift survey

J/MNRAS/246/54/  table1.dat table1.ps table1.tex  (102Kb)

Sharples R., Walker A., Cropper M. Mon. Not. R. Astron. Soc. 246, 54 (1990) Kinematics of the late M-giants in Baade’s window

J/MNRAS/246/110/ tablea1.dat (396Kb) McGilchrist M.M., Baldwin J.E., Riley J.M., Titterington D.J., Waldram E.M., Warner P.J. Mon. Not. R. Astron. Soc. 246, 110 (1990) The 7C survey of radio sources at 151 MHz - two regions centred at RA 10h 28, Dec. 41 and RA 06h 28, Dec. 45

J/MNRAS/246/169/ table3.dat table4.dat table5.dat table6.dat (20Kb) Pineault S., Chastenay P. Mon. Not. R. Astron. Soc. 246, 169 (1990) Radio continuum observations of a galactic field centred on the supernova remnant G 73.9+0.9

J/MNRAS/246/256 See VIII/22/

J/MNRAS/246/433/ table3.dat table3.tex (63Kb) Maddox S.J., Efstathiou G., Sutherland W.J. Mon. Not. R. Astron. Soc. 246, 433 (1990) The APM galaxy survey. II. Photometric corrections.

J/MNRAS/246/576/ table4.dat (25Kb) Molinari E., Buzzoni A., Chincarini G. Mon. Not. R. Astron. Soc. 246, 576 (1990) Photometry of distant clusters of galaxies: evolutionary features in the cluster 2158+0351 (z=0.45)

J/MNRAS/247/182/ table5 (56Kb) BEAUCHEMIN. M., BORRA. E.F., EDWARDS G. Mon. Not. R. Astron. Soc. 247, 182 (1990) A homogeneous catalogue of quasar candidates found with slitless spectroscopy

J/MNRAS/248/112 See VII/169/

J/MNRAS/248/398 See VIII/45/

J/MNRAS/249/218/ table1 table2 (11Kb)

Clowes R.G., Campusano L.E. Mon. Not. R. Astron. Soc. 249, 218 (1991) A 100-200 Mpc group of quasars

J/MNRAS/250/314/ table3 table4 table7 (28Kb) NOBLE R.G., DICKENS R.J., BUTTRESS J., GRIFFITHS W.K., PENNY A.J. Mon. Not. R. Astron. Soc. 250, 314 (1991) Main-sequence CCD photometry of the globular cluster omega Cen

J/MNRAS/250/683/ table3.dat (35Kb) Sagar R., Griffiths W.K. Mon. Not. R. Astron. Soc. 250, 683 (1991) CCD photometry of the distant young cluster NGC 7510.

J/MNRAS/251/46 See VIII/23/

J/MNRAS/251/253 See VIII/30/

J/MNRAS/251/330/ table2 (85Kb) WRIGHT A.E., WARK R.M., TROUP E., OTRUPCEK R., JENNINGS D., HUNT A., COOKE D.J. Mon. Not. R. Astron. Soc. 251, 330 (1991) Flux densities at 8400MHz for a large sample of radio sources

J/MNRAS/251/482/ table6 (11Kb) Boyle B.J., Jones L.R., Shanks T. Mon. Not. R. Astron. Soc. 251, 482 (1991) A spectroscopic survey of faint QSOs

J/MNRAS/252/357/ table1 table2 (76Kb) FERRARO F.R., CLEMENTINI G., FUSI PECCI F., BUONANNO R. Mon. Not. R. Astron. Soc. 252, 357 (1991) CCD-photometry of galactic globular clusters - III. NGC 6171.

J/MNRAS/253/485/ notes.dat qigc.dat (2.4 Mb) Rowan-Robinson M., Saunders W., Lawrence A., Leech K. Mon. Not. R. Astron. Soc. 253, 485 (1991) The QMW IRAS galaxy catalogue : a highly complete and reliable IRAS 60-mu-m galaxy catalogue.

J/MNRAS/253/686/ table3 table5 (20Kb) COLLESS M., ELLIS R.S., TAYLOR K., SHAW G. Mon. Not. R. Astron. Soc. 253, 686 (1991) A spectroscopic survey of faint compact objects

J/MNRAS/254/655 See VIII/60/

J/MNRAS/257/257/ app-a app-b app-c app-d app-e app-f table8 table9 (148Kb) Reid N. Mon. Not. R. Astron. Soc. 257, 257 (1992) Proper motions from Schmidt plates. II. The Hyades.

J/MNRAS/258/1 See VII/175/

J/MNRAS/259/233/ t3.tex table3 (36Kb) SODRE L., CAPELATO H.V., STEINER J.E., PROUST D., MAZURE A. Mon. Not. R. Astron. Soc. 259, 233 (1992) The cluster of galaxies SC2008-57(A3667)

J/MNRAS/259/563/ table6 table7 table8 (184Kb) Argue A.N., Bunclark P.S, Irwin M.J., Lampens P., Sinachopoulos D., Wayman P.A. Mon. Not. R. Astron. Soc. 259, 563 (1992) Double star CCD astrometry and photometry

J/MNRAS/260/77/  table1  (79Kb)

Pounds K.A., et al. Mon. Not. R. Astron. Soc. 260, 77 (1993) The ROSAT Wide Field Camera all-sky survey of extreme-ultraviolet sources - I. The Bright Source Catalogue

J/MNRAS/260/202/ table2 (12Kb) HAWKINS M.R.S., VERON P. Mon. Not. R. Astron. Soc. 260, 202 (1993) The quasar luminosity function from a variability-selected sample

J/MNRAS/260/782/ table4 (80Kb) BALONA L.A., JERZYKIEWICZ M.

Mon. Not. R. Astron. Soc. 260, 782 (1993) CCD photometry of two young Large Magellanic Cloud clusters: NGC 2004 and 2100

J/MNRAS/261/657/ table3 (66Kb) DEMERS S., IRWIN M.J Mon. Not. R. Astron. Soc. 261, 657 (1993) Deep CCD photometry of the dwarf spheroidal galaxy Leo II.

J/MNRAS/262/1057 See VIII/24/

J/MNRAS/263/25 See VIII/25/

J/MNRAS/263/425/ reftab4 table2 table3 table4a table4b (32Kb) MILLER P., RAWLINGS S., SAUNDERS R. Mon. Not. R. Astron. Soc. 263, 425 (1993) The radio and optical properties of the z 0.5 BQS quasars

J/MNRAS/264/273/ table1 (126Kb) Ferraro F.R., Clementini G., Fusi Pecci F., Vitiello E., Buonanno R. Mon. Not. R. Astron. Soc. 264, 273 (1993) On the giant, horizontal and asymptotic branches of galactic globular clusters - V. CCD photometry of NGC 1261.

J/MNRAS/265/L25 See VI/80/

J/MNRAS/266/65 See II/211/

J/MNRAS/266/317/ table1 (9Kb) CLOWES R.G., CAMPUSANO L.E. Mon. Not. R. Astron. Soc. 266, 317 (1994) Observations of quasars in ESO/SERC field 927

J/MNRAS/266/805 See VI/80/

J/MNRAS/267/711/ table1 table2 table3 table4 table5 table6a table6b table7 table8 (136Kb) WHITELOCK P., MENZIES J., FEAST M., MARANG F., CARTER B., ROBERTS G., CATCHPOLE R., CHAPMAN J. Mon. Not. R. Astron. Soc. 267, 711 (1994) High Mass-Loss AGB Stars in the South Galactic Cap

J/MNRAS/268/305/ tablea1 (20Kb) Hook I.M., McMahon R.G., Boyle B.J., Irwin M.J. Mon. Not. R. Astron. Soc. 268, 305 (1994) The variability of optically selected quasars

J/MNRAS/268/602/ pti core.dat (18Kb) Jones P.A., McAdam W.B., Reynolds J.E. Mon. Not. R. Astron. Soc. 268, 602 (1994) Compact radio cores in radio galaxies.

J/MNRAS/269/151/ table1.dat table2.dat table3.dat table4.dat (44Kb) Dalton G.B., Efstathiou G., Maddox S.J., Sutherland W.J. Mon. Not. R. Astron. Soc. 269, 151 (1994) +1995MNRAS.273..258D (Erratum) The APM galaxy survey. IV. Redshifts of rich clusters of galaxies.

J/MNRAS/269/849/ table2 (12Kb) Alpar M.A., Baykal A. Mon. Not. R. Astron. Soc. 269, 849 (1994) Expectancy of large pulsar glitches: a comparison of models with the observed glitch sample

J/MNRAS/272/41 See VI/64/

J/MNRAS/272/391/ table1 table2 table3 table4 table5 table6 table7 table8 (55Kb) FERRARO F.R., FUSI PECCI F., TESTA V., GREGGIO L., CORSI C.E., BUONANNO R., TERNDRUP D.M., ZINNECKER H. Mon. Not. R. Astron. Soc. 272, 391 (1995) Globular clusters in the Magellanic Clouds - II. IR-array photometry for 12 globular clusters and contributions to the integrated cluster light.

J/MNRAS/272/699/ table1 table2 (23Kb) BENN C.R. Mon. Not. R. Astron. Soc. 272, 699 (1995) The 5C13 deep radio survey (16h, +42)

J/MNRAS/273/157/ catalog (27Kb)

Glazebrook K., Ellis R., Colless M., Broadhurst T., Allington-Smith J., Tanvir N. Mon. Not. R. Astron. Soc. 273, 157 (1995) A faint galaxy redshift survey to B = 24.

J/MNRAS/273/411/ spectra.ps table3.dat (2.3 Mb) Lorimer D.R., Yates J.A., Lyne A.G., Gould D.M. Mon. Not. R. Astron. Soc. 273, 411 (1995) Multifrequency flux density measurements of 280 pulsars.

J/MNRAS/273/1097/ table6.dat table7.dat tablea2.dat tablea3.dat tablea4.dat tablea5.dat tablea6.dat tablea7.dat (95Kb) Jorgensen, I., Franx, M., Kjaergaard, P. Mon. Not. R. Astron. Soc. 273, 1097 (1995) Multicolour CCD surface photometry for E and S0 galaxies in 10 clusters

J/MNRAS/274/75/  table1.dat  (10Kb)

Crawford C.S., Edge A.C., Fabian A.C., Allen S.W., Bohringer H., Ebeling H., McMahon R.G., Voges W. Mon. Not. R. Astron. Soc. 274, 75 (1995) Optical spectroscopy of the ROSAT X-ray brightest clusters. II.

J/MNRAS/274/324/ tables.dat (43Kb) Green D.A., Riley J.M. Mon. Not. R. Astron. Soc. 274, 324 (1995) 408-MHz observations of two fields in the region 10h30m, +46deg. I. The surveys

J/MNRAS/274/447 See VIII/31/

J/MNRAS/274/1071/ table2 table3 table4 table5 (61Kb) Collins C.A., Guzzo L., Nichol R.C., Lumsden S.L. Mon. Not. R. Astron. Soc. 274, 1071 (1995) The Edinburgh-Durham Southern Galaxy Catalogue - VII. The Edinburgh-Milano cluster redshift survey

J/MNRAS/274/1107/ experts.dat gdevauc.dat hcorwin.dat jhuchra.dat rbuta.dat svdbergh.dat (177Kb) Naim A., Lahav O., Buta R.J., Corwin H.G.Jr., De Vaucouleurs G., Dressler A., Huchra J.P., Van Den Bergh S., Raychaudhury S., Sodre L.Jr, Storrie-Lombardi M.C. Mon. Not. R. Astron. Soc. 274, 1107 (1995) A comparative study of morphological classifications of APM galaxies

J/MNRAS/274/1165/ catalog.dat catalog.ps notes.txt (261Kb) Pye J.P., McGale P.A., Allan D.J., Barber C.R., Bertram D., Denby M., Page C.G., Ricketts M.J., Stewart B.C., West R.G. Mon. Not. R. Astron. Soc. 274, 1165 (1995) The ROSAT Wide Field Camera all-sky survey of extreme-ultraviolet sources - II. The 2RE Source Catalogue

J/MNRAS/274/1194/ appena appenb table1 table1.tex (102Kb) Mason K.O., Hassall B.J.M., Bromage G.E., Buckley D.A.H., Naylor T., O’Donoghue D., Watson M.G., Bertram D., Branduardi-Raymont G., Charles P.A., Cooke B., Elliott K.H., Hawkins M.R.S., Hodgkin S.T., Jewell S.J., Jomaron C.M., Sekiguchi K., Kellett B.J., Lawrence A., McHardy I., Mittaz J.P.D., Pike C.D., Ponman T.J., Schmitt J., Voges W., Wargau W., Wonnacott D. Mon. Not. R. Astron. Soc. 274, 1194 (1995) Optical identification of EUV sources from the ROSAT Wide Field Camera all-sky survey.

J/MNRAS/275/169 See II/211/

J/MNRAS/275/591/ main obs (171Kb) IBATA R.A., GILMORE G. Mon. Not. R. Astron. Soc. 275, 591 (1995) The Outer Regions of the Galactic Bulge: I. Observations.

J/MNRAS/275/1102/ table1a table1b table1c (24Kb) Hawkins M.R.S., Veron P. Mon. Not. R. Astron. Soc. 275, 1102 (1995) The evolution of the quasar luminosity function

J/MNRAS/276/33/  tables.dat  (253Kb)

Boyle B.J., Shanks T., Croom S.M. Mon. Not. R. Astron. Soc. 276, 33 (1995) Faint UBVRI CCD sequences for wide-field surveys. I.

J/MNRAS/276/57/  table1 table2 table3 table5  (48Kb)

Codella C., Palumbo G.G.C., Pareschi G., Scappini F., Caselli P., Attolini M.R. Mon. Not. R. Astron. Soc. 276, 57 (1995)

IRAS-selected Galactic star-forming regions. - II. Water maser detections in the extended sample

J/MNRAS/276/614/ table2 table3 table3.tex (81Kb) Lacy M., Riley J.M., Waldram E.M., McMahon R.G., Warner P.J. Mon. Not. R. Astron. Soc. 276, 614 (1995) A radio-optical survey of the North Ecliptic Cap.

J/MNRAS/276/689/ table4 (31Kb) Ettori S., Guzzo L., Tarenghi M. Mon. Not. R. Astron. Soc. 276, 689 (1995) A study of the large-scale distribution of galaxies in the South Galactic Pole region - I. The data.

J/MNRAS/277/125/ table6a table6b (260Kb) Hau G.K.T., Ferguson H.C., Lahav O., Lynden-Bell D. Mon. Not. R. Astron. Soc. 277, 125 (1995) Visual search for galaxies near the northern crossing of the Supergalactic Plane by the Milky Way.

J/MNRAS/277/1477/ catalog (37Kb) Padovani P., Giommi P. Mon. Not. R. Astron. Soc. 277, 1477 (1995) A sample-oriented catalogue of BL Lacertae objects

J/MNRAS/278/1025/ fields.dat galaxy.dat holes.dat (1.1 Mb) Loveday J. Mon. Not. R. Astron. Soc. 278, 1025 (1996) The APM Bright Galaxy Catalogue

J/MNRAS/279/95 See VI/80/

J/MNRAS/279/459/ table1 tablea1 tables.tex (47Kb) Clements D.L., Sutherland W.J., Saunders W., Efstathiou G.P., McMahon R.G., Maddox S., Lawrence A., Rowan-Robinson M. Mon. Not. R. Astron. Soc. 279, 459 (1996) A new large sample of ultraluminous IRAS galaxies.

J/MNRAS/279/595/ table9 table9.ps (839Kb) Drinkwater M.J., Currie M.J., Young C.K., Hardy E., Yearsley J.M. Mon. Not. R. Astron. Soc. 279, 595 (1996) Blue compact dwarf galaxies and new velocities in Virgo.

J/MNRAS/281/799/ refs.dat table2.dat table34.dat (52Kb) Ebeling H., Voges W., Bohringer H., Edge A.C., Huchra J.P., Briel U.G. Mon. Not. R. Astron. Soc. 281, 799 + Erratum vol. 283, p. 1103 (1996) Properties of the X-ray-brightest Abell-type clusters of galaxies (XBACs) from ROSAT All-Sky Survey Data - I. The sample.

J/MNRAS/282/779/ 7c.dat (425Kb) Waldram E.M., Yates J.A., Riley J.M., Warner P.J. Mon. Not. R. Astron. Soc. 282, 779 (1996) The 7C Survey of Radio Sources at 151 MHz - a region covering RA 9h to 16h and Dec 20 deg to 35 deg.

J/MNRAS/283/367/ table3 (43Kb) di Nella H., Couch W.J., Paturel G., Parker Q.A. Mon. Not. R. Astron. Soc. 283, 367 (1996) Are the Perseus-Pisces chain and the Pavo-Indus wall connected?

J/MNRAS/284/85 See VIII/34/

J/MNRAS/284/265/ catalog.dat notes.dat (61Kb) Howarth I.D., Siebert K.W., Hussain G.A.J., Prinja R.K. Mon. Not. R. Astron. Soc. 284, 265 (1997) Cross-correlation characteristics of OB stars from IUE spectroscopy.

J/MNRAS/285/1 See I/241/

J/MNRAS/285/898/ table3 table4 (19Kb) Venturi T., Bardelli S., Morganti R., Hunstead R.W. Mon. Not. R. Astron. Soc. 285, 898 (1997) Radio properties of the Shapley Concentration - I. The Abell cluster A3556

J/MNRAS/286/500/ absolute.dat relative.dat stars.dat (149Kb) Malyuto V., Oestreicher M.O., Schmidt-Kaler Th. Mon. Not. R. Astron. Soc. 286, 500 (1997)

Quantitative Spectral Classification of Galactic Disk K-M Stars from Spectrophotometric Measurements

J/MNRAS/287/293/ tab2-8 table1 table9 tables.tex (102Kb) Maoz D., Ofek E.O., Shemi A. Mon. Not. R. Astron. Soc. 287, 293 (1997) Evidence for a new class of extreme ultraviolet sources

J/MNRAS/287/472/ table3.dat (27Kb) Di Nella H., Couch W.J., Parker Q.A., Paturel G. Mon. Not. R. Astron. Soc. 287, 472 (1997) Unveiling a connection between large-scale structures behind the southern Milky Way

J/MNRAS/287/867/ note1.dat note2.dat table1.dat table2.dat (94Kb) Kilkenny D., O’Donoghue D., Koen C., Stobie R.S., Chen A. Mon. Not. R. Astron. Soc. 287, 867 (1997) The Edinburgh-Cape Blue Object Survey - II. Zone 1 - the North Galactic Cap

J/MNRAS/289/225/ table1 (35Kb) Staveley-Smith L., Sault R.J., Hatzidimitriou D., Kesteven M.J., McConnell D. Mon. Not. R. Astron. Soc. 289, 225 (1997) An H I aperture synthesis mosaic of the Small Magellanic Cloud

J/MNRAS/289/263/ catalog (55Kb) Dalton G.B., Maddox S.J., Sutherland W.J., Efstahiou G. Mon. Not. R. Astron. Soc. 289, 263 (1997) The APM galaxy survey - V. Catalogues of galaxy clusters.

J/MNRAS/289/700 See VI/89A/

J/MNRAS/289/729/ bvri.dat jhk.dat (31Kb) Oestreicher M.O., Schmidt-Kaler T., Wargau W. Mon. Not. R. Astron. Soc. 289, 729 (1997) Red supergiants in the LMC. I. BVRIJHK photometry, magnitudes and intrinsic colours.

J/MNRAS/289/745/ table3.dat (9Kb) Rosino L., Ortolani S., Barbuy B., Bica E. Mon. Not. R. Astron. Soc. 289, 745 (1997) NGC 6749: a metal-poor halo globular cluster in a disc field rich in Mira variables.

J/MNRAS/291/261/ table1.dat table2.dat tables.tex (66Kb) Walsh A.J., Hyland A.R., Robinson G., Burton M.G Mon. Not. R. Astron. Soc. 291, 261 (1997) Studies of ultracompact H II regions. I. Methanol maser survey of IRAS-selected sources.

J/MNRAS/292/419/ refs.dat table1.dat table1.tex table2.dat table2.tex table3.dat table3.tex table4.dat table4.tex table5.dat table5.tex table6.dat table6.tex (440Kb) White D.A., Jones C., Forman W. Mon. Not. R. Astron. Soc., 292, 419-467 (1997) An investigation of cooling flows and general cluster properties from an X-ray image deprojection analysis of 207 clusters of galaxies.

J/MNRAS/292/499/ notes.dat table4.dat (148Kb) Saglia R.P., Burstein D., Baggley G., Bertschinger E., Colless M.M., Davies R.L., McMahan R.K.Jr, Wegner G. Mon. Not. R. Astron. Soc. 292, 499 (1997) The peculiar motions of early-type galaxies in two distant regions - III. The photometric data.

J/MNRAS/293/257 See VIII/60/

J/MNRAS/294/607/ catalog.dat (514Kb) Vessey S.J., Green D.A. Mon. Not. R. Astron. Soc. 294, 607 (1998) The 7C(G) survey of radio sources at 151MHz - the Galactic plane at 80deg l 104deg and 126deg l 180deg, for —b— 5.5deg.

J/MNRAS/296/599/ table2.dat (15Kb) Bardelli S., Zucca E., Zamorani G., Vettolani G., Scaramella R. Mon. Not. R. Astron. Soc. 296, 599 (1998) A study of the core of the Shapley concentration. III. Properties of the clusters in the A3558 complex.

J/MNRAS/296/839/ pdf.dat (145Kb) Hopkins A.M., Mobasher, B. Cram, L., Rowan-Robinson, M. Mon. Not. R. Astron. Soc. 296, 839 (1998)

The Phoenix Deep Survey: 1.4-GHz source counts

J/MNRAS/297/255/ catalog.dat (110Kb) Knapen J.H. Mon. Not. R. Astron. Soc. 297, 255 (1998) Statistical properties of HII regions in the disc of M100

J/MNRAS/297/885/ table2.dat (39Kb) Glazebrook K., Abraham R., Santiago B., Ellis R., Griffiths R. Mon. Not. R. Astron. Soc. 297, 885 (1998) The physical parameters of the evolving population of faint galaxies

J/MNRAS/298/332/ tablea1.dat (46Kb) Rocha-Pinto H.J., Maciel W.J. Mon. Not. R. Astron. Soc. 298, 332 (1998) Metallicity effects on the chromospheric activity-age relation for late-type dwarfs

J/MNRAS/298/637/ 100553.dat 104257.dat 133238.dat (182Kb) Pooley D.M., Waldram E.M., Riley J.M. Mon. Not. R. Astron. Soc. 298, 637 (1998) The 7C survey of radio sources at 151 MHz: three regions centred at RA=10h05min, DE=53deg54’, RA=10h42min, DE=57deg36’ and RA=13h32min, DE=38deg10’

J/MNRAS/298/692/ fits/* list.dat (241.3 Mb) Fender R.P., Southwell K.A., Tzioumis A.K. Mon. Not. R. Astron. Soc. 298, 692 (1998) A radio survey of supersoft, persistent and transient X-ray sources in the Magellanic Cloud

J/MNRAS/298/753/ photom.dat table4.dat (486Kb) Pigulski A., Kolaczkowski Z. Mon. Not. R. Astron. Soc. 298, 753 (1998) Mapping the βCephei instability strip: photometric variability of stars in the central part of the Cygnus OB2 association.

J/MNRAS/299/1/  table3.dat table4.dat table5.dat table6.dat tables.tex  (855Kb)

Sagar R., Griffiths W.K. Mon. Not. R. Astron. Soc. 299, 1 (1998) BVI CCD photometry of the distant open star clusters Berkeley 81, Berkeley 99, NGC 6603 and NGC 7044.

J/MNRAS/299/834/ table.dat (96Kb) Tosi M., Pulone L., Marconi G., Bragaglia A. Mon. Not. R. Astron. Soc. 299, 834 (1998) Old open clusters : the interesting case of Berkeley 21

J/MNRAS/300/303/ table1.dat (15Kb) Treyer M.A., Ellis R.S., Milliard B., Donas J., Bridges R.J. Mon. Not. R. Astron. Soc. 300, 303 (1998) An ultraviolet-selected galaxy redshift survey: new estimates of the local star formation rate.

J/MNRAS/300/417/ table5.dat (191Kb) Ratcliffe A., Shanks T., Parker Q.A., Broadbent A., Watson F.G., Oates A.P., Collins C.A., Fong R. Mon. Not. R. Astron. Soc. 300, 417 (1998) The Durham/UKST galaxy redshift survey - V. The catalogue.

J/MNRAS/300/665/ n1754.dat n1835.dat n1898.dat n1916.dat n2005.dat n2019.dat (4.2 Mb) Olsen K.A.G., Hodge P.W., Mateo M., Olszewski E.W., Schommer R.A., Suntzeff N.B., Walker A.R. Mon. Not. R. Astron. Soc., 300, 665 (1998) HST colour-magnitude diagrams of six old globular clusters in the LMC

J/MNRAS/300/790 See VIII/60/

J/MNRAS/301/203/ 67p.dat 68p.dat 69p.dat 70p.dat (91Kb) Riley J.M., Green D.A. Mon. Not. R. Astron. Soc. 301, 203 (1998) A 408-MHz variability study of two fields in the region 06:30 +45^{∘}

J/MNRAS/301/640/ table2.dat (130Kb) Walsh A.J., Burton M.G., Hyland A.R., Robinson G. Mon. Not. R. Astron. Soc. 301, 640 (1998) Studies of ultracompact H II regions - II. High-resolution radio continuum and methanol maser survey.

J/MNRAS/301/881/ refs.dat table3.dat (35Kb) Ebeling H., Edge A.C., Boehringer H., Allen S.W., Crawford C.S., Fabian A.C., Voges W., Huchra J.P.

Mon. Not. R. Astron. Soc. 301, 881 (1998) The ROSAT brightest cluster sample - I. The compilation of the sample and the cluster log N-log S distribution.

J/MNRAS/301/1031/ lhs2065.dat lp944-20.dat mucol.dat vb8.dat (1.3 Mb) Tinney C.G., Reid I.N. Mon. Not. R. Astron. Soc. 301, 1031 (1998) High resolution spectra of Very Low-Mass Stars

J/MNRAS/302/222/ table3.dat (197Kb) Ciliegi P., McMahon R.G., Miley G., Gruppioni C., Rowan-Robinson M., Cesarsky C., Danese L., Franceschini A., Genzel R., Lawrence A., Lemke D., Oliver S., Puget J.-L., Rocca-Volmerange B. Mon. Not. R. Astron. Soc. 302, 222 (1999) A deep VLA survey at 20 cm of the ISO ELAIS survey regions.

J/MNRAS/304/297/ catalog.dat (144Kb) Morshidi-Esslinger Z., Davies J.I., Smith R.M. Mon. Not. R. Astron. Soc. 304, 297 (1999) An automated search for nearby low-surface-brightness galaxies. I. The catalogue

J/MNRAS/304/733/ ( total of 7 files) (6.7 Mb) Petterson O.K.L., Tobin W. Mon. Not. R. Astron. Soc., 304, 733 (1999) Beta Pictoris: the variable CaII H + K absorptions 1994-1996

J/MNRAS/305/259/ table3.dat table5.dat table7.dat tables.ps tables.tex (774Kb) Wegner G., Colless M., Saglia R.P., McMahan R.K.Jr, Davies R.L., Burstein D., Baggley G. Mon. Not. R. Astron. Soc. 305, 259 (1999) The peculiar motions of early-type galaxies in two distant regions - II. The spectroscopic data.

J/MNRAS/305/297/ catalog.dat (107Kb) Gruppioni C., Ciliegi P., Rowan-Robinson M., Cram L., Hopkins A., Cesarsky C., Danese L., Franceschini A., Genzel R., Lawrence A., Lemke D., McMahon R.G., Miley G., Oliver S., Puget J.-L., Rocca-Volmerange B. Mon. Not. R. Astron. Soc. 305, 297 (1999) A 1.4-GHz survey of the southern European Large-Area ISO Survey region.

J/MNRAS/306/279/ notes.dat table1.dat table2.dat table3.dat (94Kb) Schmidt-Kaler T., Gochermann J., Oestreicher M.O., Grothues H.-G., Tappert C., Zaum A., Berghoefer T., Brugger H.R. Mon. Not. R. Astron. Soc. 306, 279 (1999) UBV photometry of Galactic foreground and LMC member stars - III. LMC member stars - a new data base.

J/MNRAS/306/361/ table1.dat (105Kb) Sung H., Bessell M.S. Mon. Not. R. Astron. Soc. 306, 361 (1999) UBVI CCD photometry of M 35 (NGC 2168).

J/MNRAS/306/592/ tables.dat (145Kb) Croom S.M., Ratcliffe A., Parker Q.A., Shanks T., Boyle B.J., Smith R.J. Mon. Not. R. Astron. Soc. 306, 592 (1999) Faint UBVRI CCD sequences for wide-field surveys. II. UBVR sequences at δ=-30^{∘}

J/MNRAS/306/708/ table1.dat (37Kb) Georgakakis A., Mobasher B., Cram L., Hopkins A., Lidman C., Rowan-Robinson M. Mon. Not. R. Astron. Soc. 306, 708 (1999) The Phoenix Survey: optical and near-infrared observations of faint radio sources.

J/MNRAS/307/236/ table1.dat table2.dat (21Kb) Chamaraux P., Masnou J.-L., Kazes I., Saito M., Takata T., Yamada T. Mon. Not. R. Astron. Soc. 307, 236 (1999) H I observations of galaxies behind the Milky Way in the Puppis region.

J/MNRAS/308/473/ sp/* table1.dat (235Kb) Fender R.P., Hanson M.M., Pooley G.G. Mon. Not. R. Astron. Soc. 308, 473 (1999) Infrared spectroscopic variability of Cygnus X-3 in outburst and quiescence.

J/MNRAS/308/897/ f15.dat ngw.dat qdot93.dat qdotv.dat sgw.dat (579Kb) Lawrence A., Rowan-Robinson M., Ellis R.S., Frenk C.S., Efstathiou G., Kaiser N., Saunders W., Parry I.R., Xia X., Crawford J. Mon. Not. R. Astron. Soc. 308, 897 (1999) The QDOT all-sky IRAS galaxy redshift survey.

J/MNRAS/310/210/ hrd con5.mpg hrd con6.mpg hrd con7.mpg hrd dec5.mpg hrd dec6.mpg hrd dec7.mpg hrd inc5.mpg hrd inc6.mpg hrd inc7.mpg hrd inst.mpg models.dat notes.dat table1.dat (20.8 Mb) Shevchenko V.S., Ezhkova O.V., Ibrahimov M.A., van den Ancker M.E., Tjin A Djie H.R.E. Mon. Not. R. Astron. Soc. 310, 210 (1999) The stellar composition of the star formation region CMa R1. I. Results from new photometric and spectroscopic classifications

J/MNRAS/311/456/ rixos.dat table1.dat table3.dat (95Kb) Mason K.O., Carrera F.J., Hasinger G., Andernach H., Aragon-Salamanca A., Barcons X., Bower R., Brandt W.N., Branduardi-Raymont G., Burgos-Martin J., Cabrera-Guerra F., Carballo R., Castander F., Ellis R.S., Gonzalez-Serrano J.I., Martinez-Gonzalez E., Martin-Mirones J.M., McMahon R.G., Mittaz J.P.D., Nicholson K.L., Page M.J., Perez-Fournon I., Puchnarewicz E.M., Romero-Colmenero E., Schwope A.D., Vila B., Watson M.G., Wonnacott D. Mon. Not. R. Astron. Soc. 311, 456 (2000) The ROSAT International X-ray/Optical Survey (RIXOS): source catalogue

J/MNRAS/311/741/ table1.dat table2.dat (28Kb) Murphy M.T., Bessell M.S. Mon. Not. R. Astron. Soc. 311, 741 (2000) Catalogue of candidate emission-line objects in the Small Magellanic Cloud.

J/MNRAS/312/442/ table7.dat (39Kb) Sullivan M., Treyer M.A., Ellis R.S., Bridges T.J., Milliard B., Donas J. Mon. Not. R. Astron. Soc. 312, 442 (2000) An ultraviolet-selected galaxy redshift survey. II. The physical nature of star formation in an enlarged sample.

J/MNRAS/312/540/ table2.dat (25Kb) Bardelli S., Zucca E., Zamorani G., Moscardini L., Scaramella R. Mon. Not. R. Astron. Soc. 312, 540 (2000) A study of the core of the Shapley Concentration - IV. Distribution of intercluster galaxies and supercluster properties.

J/MNRAS/312/557/ table1.dat (37Kb) Loveday J. Mon. Not. R. Astron. Soc. 312, 557 (2000) The K-band Luminosity Function of Nearby Field Galaxies

J/MNRAS/313/43/  tablea1.dat  (98Kb)

Hoogerwerf R. Mon. Not. R. Astron. Soc. 313, 43 (2000) OB association members in the ACT and TRC catalogues.

J/MNRAS/313/347/ tablea2.dat tablea3.dat tablea4.dat tablea5.dat (35Kb) Pinfield D.J., Hodgkin S.T., Jameson R.F., Cossburn M.R., Hambly N.C., Devereux N. Mon. Not. R. Astron. Soc. 313, 347 (2000) A six-square-degree survey for Pleiades low-mass stars and brown dwarfs.

J/MNRAS/313/469/ table3.dat (92Kb) Smith R.J., Lucey J.R., Hudson M.J., Schlegel D.J., Davies R.L. Mon. Not. R. Astron. Soc. 313, 469 (2000) Streaming motions of galaxy clusters within 12000km/s I. New spectroscopic data.

J/MNRAS/313/641/ table2.dat table4.dat table6.dat (119Kb) Brosch N., Brook A., Wisotzki L., Almoznino E., Prialnik D., Bowyer S., Lampton M. Mon. Not. R. Astron. Soc. 313, 641 (2000) FAUST observations in the Fourth Galactic Quadrant.

J/MNRAS/315/839/ table1.dat table7.dat table9.dat (38Kb) Minns A.R., Riley J.M. Mon. Not. R. Astron. Soc. 315, 839 (2000) Compact radio sources variable at 151MHz - I. Catalogue of sources.

J/MNRAS/316/326/ tablec1.dat (23Kb) Hasegawa T., Wakamatsu K.-I., Malkan M., Sekiguchi K., Menzies J.W., Parker Q.A., Jugaku J., Karoji H., Okamura S. Mon. Not. R. Astron. Soc. 316, 326 (2000) Large-scale structure of galaxies in the Ophiuchus region.

J/MNRAS/316/514/ table4a.dat table4b.dat (79Kb) Adelman S.J., Caliskan H., Kocer D., Hamami I., Tektunali H.G. Mon. Not. R. Astron. Soc. 316, 514 (2000)

Elemental abundance analyses with DAO spectrograms. XXIII. The superficially normal stars 28 And (A7 III) and 99 Her (F7 V).

J/MNRAS/317/55 See VII/221/

J/MNRAS/317/234/ tables.dat (36Kb) Perez-Ramirez D., Knapen J.H., Peletier R.F., Laine S., Doyon R., Nadeau D. Mon. Not. R. Astron. Soc. 317, 234 (2000) Circumnuclear regions in barred spiral galaxies I. Near-infrared imaging

J/MNRAS/317/460/ refs.dat stars.dat table1.dat tablea2.dat (107Kb) Feast M.W., Whitelock P.A. Mon. Not. R. Astron. Soc. 317, 460 (2000) Mira kinematics from Hipparcos data: a Galactic bar to beyond the Solar circle.

J/MNRAS/317/801/ table1.dat table3.dat (29Kb) Waddington I., Windhorst R.A., Dunlop J.S., Koo D.C., Peacock J.A. Mon. Not. R. Astron. Soc. 317, 801 (2000) The LBDS Hercules sample of mJy radio sources at 1.4GHz - I. Multicolour photometry.

J/MNRAS/318/333/ refs.dat table1.dat (21Kb) Ebeling H., Edge A.C., Allen S.W., Crawford C.S., Fabian A.C., Huchra J.P. Mon. Not. R. Astron. Soc. 318, 333 (2000) The ROSAT Brightest Cluster Sample - IV. The extended sample.

J/MNRAS/318/703/ table2.dat (130Kb) Balogh M.L., Morris S.L. Mon. Not. R. Astron. Soc. 318, 703 (2000) H-alpha photometry of Abell 2390

J/MNRAS/318/1073/ table2.dat (25Kb) Page M.J., Mittaz J.P.D., Carrera F.J. Mon. Not. R. Astron. Soc. 318, 1073 (2000) A survey of hard spectrum ROSAT sources - I. X-ray source catalogue.

J/MNRAS/318/1206/ table2.dat (19Kb) Cooke J.A., Neill Reid I. Mon. Not. R. Astron. Soc. 318, 1206 (2000) High proper motion stars in Kapteyn Selected Area 94.

J/MNRAS/319/215/ tablea1.dat (51Kb) Han Z., Tout C.A., Eggleton P.P. Mon. Not. R. Astron. Soc. 319, 215 (2000) Low and intermediate-mass close binary evolution and the initial-final mass relation

J/MNRAS/320/387/ table1.dat table2.dat table4.dat (33Kb) Bardelli S., Zucca E., Baldi A. Mon. Not. R. Astron. Soc. 320, 387 (2001) A study of the core of the Shapley Concentration - V. The A3528 complex: a young merger event?

J/MNRAS/320/451/ table1.dat table2.dat table4.dat table5.dat table6.dat (46Kb) Beers T.C., Rossi S., O’Donoghue D., Kilkenny D., Stobie R.S., Koen C., Wilhelm R. Mon. Not. R. Astron. Soc. 320, 451 (2001) Metallicity estimates for A-, F-, and G-type stars from the Edinburgh-Cape Blue Object Survey.

J/MNRAS/321/77/  table1.dat table2.dat table3.dat table4.dat table5.dat table6.dat  (48Kb)

Glass I.S., Matsumoto S., Carter B.S., Sekiguchi K. Mon. Not. R. Astron. Soc. 321, 77 (2001) Large-amplitude variables near the Galactic Centre.

J/MNRAS/321/277/ table1.dat table2.dat (244Kb) Colless M.M., Saglia R.P., Burstein D., Davies R.L., McMahan R.K., Wegner G. Mon. Not. R. Astron. Soc. 321, 277 (2001) The peculiar motions of early-type galaxies in two distant regions - VII. Peculiar velocities and bulk motions

J/MNRAS/322/L29 See VII/223/

J/MNRAS/323/757/ notes.dat table4.dat table5.dat (42Kb) Landt H., Padovani P., Perlman E.S., Giommi P., Bignall H., Tzioumis A. Mon. Not. R. Astron. Soc. 323, 757 (2001) The Deep X-Ray Radio Blazar Survey (DXRBS) - II. New identifications.

J/MNRAS/323/839/ table1.dat (14Kb)

Smail I., Kuntschner H., Kodama T., Smith G.P., Packham C., Fruchter A.S., Hook R.N. Mon. Not. R. Astron. Soc. 323, 839 (2001) A photometric study of the ages and metallicities of early-type galaxies in A 2218.

J/MNRAS/324/33/  ab8rv.dat br22rv.dat br31rv.dat br32rv.dat wr9rv.dat wr30arv.dat  (14Kb)

Bartzakos P., Moffat A.F.J., Niemela V.S. Mon. Not. R. Astron. Soc. 324, 33 (2001) Magellanic Cloud WC/WO Wolf-Rayet stars - II. Colliding winds in binaries

J/MNRAS/324/580/ gnotes.dat snotes.dat table1.dat table2.dat table3.dat (61Kb) Daniels J., Brosch N., Almoznino E., Shemmer O., Bowyer S., Lampton M. Mon. Not. R. Astron. Soc. 324, 580 (2001) FAUST observations of ultraviolet sources in the directions of NGC 4038-39 and 6752.

J/MNRAS/324/1131/ notes.dat table3.dat table4.dat (29Kb) Venturi T., Bardelli S., Zambelli G., Morganti R., Hunstead R.W. Mon. Not. R. Astron. Soc. 324, 1131 (2001) Radio properties of the Shapley Concentration - IV. The A3528 cluster complex.

J/MNRAS/325/931/ table1.dat (32Kb) Reid I.N., Wyk F.V., Marang F., Roberts G., Kilkenny D., Mahoney S. Mon. Not. R. Astron. Soc. 325, 931 (2001) A search for previously unrecognized metal-poor subdwarfs in the Hipparcos astrometric catalogue.

J/MNRAS/325/1002/ table1.dat table2.dat (55Kb) Willis, J.P, Hewett, P.C., Warren, S.J. Mon. Not. R. Astron. Soc. 325, 1002 (2001) Luminous early-type galaxies at redshifts z 0.4. I. Observations and redshift catalogue of 581 galaxies.

J/MNRAS/325/1173/ catalog.dat (48Kb) Lari C., Pozzi F., Gruppioni C., Aussel H., Ciliegi P., Danese L., Franceschini A., Oliver S., Rowan-Robinson M., Serjeant S. Mon. Not. R. Astron. Soc. 325, 1173 (2001) A new method for ISOCAM data reduction - I. Application to the European Large Area ISO Survey Southern Field: method and results.

J/MNRAS/325/1365/ table3.dat (14Kb) Haywood M. Mon. Not. R. Astron. Soc. 325, 1365 (2001) A revision of the solar neighbourhood metallicity distribution.

J/MNRAS/325/1411/ table3.dat table4.dat (34Kb) Kedziora-Chudczer L.L., Jauncey D.L., Wieringa M.H., Tzioumis A.K., Reynolds J.E. Mon. Not. R. Astron. Soc. 325, 1411 (2001) The ATCA intraday variability survey of extragalactic radio sources.

J/MNRAS/326/473/ galaxies.dat kinem1.dat kinem2.dat (82Kb) Halliday C., Davies R. L., Kuntschner H., Birkinshaw M., Bender R., Saglia R.P., Baggley G. Mon. Not. R. Astron. Soc., 326, 473 (2001) Line-of-sight velocity distributions of low-luminosity elliptical galaxies

J/MNRAS/326/959/ table.dat (166Kb) Cenarro, A.J., Cardiel, N., Gorgas, J., Peletier, R.F., Vazdekis, A., Prada, F. Mon. Not. R. Astron. Soc. 326, 959 (2001) Empirical calibration of the near-IR CaII triplet. I. The stellar library and index definition.

J/MNRAS/326/1161/ table4.dat table23.dat (47Kb) Morley J.E., Briggs K.R., Pye J.P., Favata F., Micela G., Sciortino S. Mon. Not. R. Astron. Soc. 326, 1161 (2001) A ROSAT medium-sensitivity Galactic plane survey at 180^{∘}l 280^{∘}.

J/MNRAS/326/1455/ table1.dat (31Kb) Marcha M.J., Caccianiga A., Browne I.W.A., Jackson N. Mon. Not. R. Astron. Soc. 326, 1455 (2001) The CLASS blazar survey - I. Selection criteria and radio properties.

J/MNRAS/327/23/ table1.dat table4.dat table5.dat table6.dat table7.dat table8.dat table9.dat table10.dat (183Kb) Sagar R., Munari U., de Boer K.S. Mon. Not. R. Astron. Soc. 327, 23 (2001) A multicolour CCD photometric and mass function study of the distant southern open star clusters NGC

3105, NGC 3603, Melotte 105, Hogg 15, NGC 4815, Pismis 20 and NGC 6253.

J/MNRAS/327/249/ table4.dat (73Kb) Smith R.J., Lucey J.R., Schlegel D.J., Hudson M.J., Baggley G., Davies R.L. Mon. Not. R. Astron. Soc. 327, 249 (2001) Streaming motions of galaxy clusters within 12000km/s. II. New photometric data for the Fundamental Plane.

J/MNRAS/327/265/ table6.dat table7.dat (94Kb) Hudson M.J., Lucey J.R., Smith R.J., Schlegel D.J., Davies R.L. Mon. Not. R. Astron. Soc. 327, 265 (2001) Streaming motions of galaxy clusters within 12000km/s. III. A standardized catalogue of Fundamental Plane data.

J/MNRAS/327/588/ table1.dat (6Kb) Pimbblet K.A., Smail I., Edge A.C., Couch W.J., O’Hely E., Zabludoff A.I. Mon. Not. R. Astron. Soc. 327, 588 (2001) The Las Campanas/AAT Rich Cluster Survey - I. Precision and reliability of the photometric catalogue.

J/MNRAS/327/601/ table1.dat (9Kb) Eyer L., Wozniak P.R. Mon. Not. R. Astron. Soc. 327, 601 (2001) Photometric detection of high proper motions in dense stellar fields using Difference Image Analysis.

J/MNRAS/327/771/ table2.dat (17Kb) Fiore F., Giommi P., Vignali C., Comastri A., Matt G., Perola G.C., La Franca F., Molendi S., Tamburelli F., Antonelli L.A. Mon. Not. R. Astron. Soc. 327, 771 (2001) The BeppoSAX High Energy Large Area Survey (HELLAS). II. Number counts and X-ray spectral properties.

J/MNRAS/327/918/ table5.dat table6.dat table7.dat tablea1.dat tablea2.dat tablea3.dat (22Kb) Tolstoy E., Irwin M.J., Cole A.A., Pasquini L., Gilmozzi R., Gallagher J.S. Mon. Not. R. Astron. Soc. 327, 918 (2001) Using the Ca II triplet to trace abundance variations in individual red giant branch stars in three nearby galaxies.

J/MNRAS/327/1004/ table1.dat (34Kb) Blakeslee J.P., Lucey J.R., Barris B.J., Hudson M.J., Tonry J.L. Mon. Not. R. Astron. Soc. 327, 1004 (2001) A synthesis of data from fundamental plane and surface brightness fluctuation surveys.

J/MNRAS/327/1116/ table2.dat (14Kb) Larsen S.S., Forbes D.A., Brodie J.P. Mon. Not. R. Astron. Soc. 327, 1116 (2001) Hubble Space Telescope photometry of globular clusters in the Sombrero galaxy.

J/MNRAS/328/17/  table4.dat table5.dat table7.dat  (35Kb)

Manchester R.N., Lyne A.G., Camilo F., Bell J.F., Kaspi V.M., D’Amico N., McKay N.P.F., Crawford F., Stairs I.H., Possenti A., Kramer M., Sheppard D.C. Mon. Not. R. Astron. Soc. 328, 17 (2001) The Parkes Multi-Beam Pulsar Survey. I. Observing and data analysis systems, discovery and timing of 100 pulsars.

J/MNRAS/328/45/  binaries.dat singles.dat  (161Kb)

Montes D., Lopez-Santiago J., Galvez M.C., Fernandez-Figueroa M.J., De Castro E., Cornide M. Mon. Not. R. Astron. Soc. 328, 45 (2001) Late-type stars members of young stellar kinematic groups. I. Single stars (2001MNRAS.328...45M)

J/MNRAS/328/150/ table1.dat table2.dat table3.dat table4.dat (16Kb) Croom S.M., Warren S.J., Glazebrook K. Mon. Not. R. Astron. Soc. 328, 150 (2001) A small-area faint KX redshift survey for QSOs in the ESO Imaging Survey Chandra Deep Field South.

J/MNRAS/328/461/ table3.dat (27Kb) O’Sullivan E., Forbes D.A., Ponman T.J. Mon. Not. R. Astron. Soc. 328, 461 (2001) A catalogue and analysis of X-ray luminosities of early-type galaxies.

J/MNRAS/328/805/ table1.dat (11Kb) Outram P.J., Smith R.J., Shanks T., Boyle B.J., Croom S.M., Loaring N.S., Miller L. Mon. Not. R. Astron. Soc. 328, 805 (2001) The 2dF QSO Redshift Survey - VIII. Absorption systems in the 10k catalogue.

J/MNRAS/328/855/ table1.dat (15Kb) Wang N., Manchester R.N., Zhang J., Wu X.J., Yusup A., Lyne A.G., Cheng K.S., Chen M.Z. Mon. Not. R. Astron. Soc. 328, 855 (2001) Pulsar timing at Urumqi Astronomical Observatory: observing system and results.

J/MNRAS/328/1039 See VII/226/

J/MNRAS/328/1181/ table1.dat table2.dat table3.dat (35Kb) Navasardyan H., Petrosian A.R., Turatto M., Cappellaro E., Boulesteix J. Mon. Not. R. Astron. Soc. 328, 1181 (2001) Supernovae in isolated galaxies, in pairs and in groups of galaxies.

J/MNRAS/329/87/  table1.dat  (84Kb)

de Propris R., Couch W.J., Colless M., Dalton G.B., Collins C., Baugh C.M., Bland-Hawthorn J., Bridges T., Cannon R., Cole S., Cross N., Deeley K., Driver S.P., Efstathiou G., Ellis R.S., Frenk C.S., Glazebrook K., Jackson C., Lahav O., Lewis I., Lumsden S., Maddox S., Madgwick D., Moody S., Norberg P., Peacock J.A., Percival W., Peterson B.A., Sutherland W., Taylor K. Mon. Not. R. Astron. Soc. 329, 87 (2002) The 2dF Galaxy Redshift Survey: a targeted study of catalogued clusters of galaxies.

J/MNRAS/329/227/ notes.dat table1.dat table1b.dat table5.dat table6.dat (105Kb) Sadler E.M., Jackson C.A., Cannon R.D., McIntyre V.J., Murphy T., Bland-Hawthorn J., Bridges T., Cole S., Colless M., Collins C., Couch W., Dalton G., de Propris R., Driver S.P., Efstathiou G., Ellis R.S., Frenk C.S., Glazebrook K., Lahav O., Lewis I., Lumsden S., Maddox S., Madgwick D., Norberg P., Peacock J.A., Peterson B.A., Sutherland W., Taylor K. Mon. Not. R. Astron. Soc., 329, 227 (2002) Radio sources in the 2dF Galaxy Redshift Survey - II. Local radio luminosity functions for AGN and starforming galaxies at 1.4GHz

J/MNRAS/329/700/ list1.dat list2.dat list3.dat (176Kb) Snellen I.A.G., McMahon R.G., Hook I.M., Browne I.W.A. Mon. Not. R. Astron. Soc. 329, 700 (2002) Automated optical identification of a large complete northern hemisphere sample of flat-spectrum radio sources with S_{6cm} 200mJy

J/MNRAS/329/775/ table4.dat table5.dat (12Kb) Roy S., Rao A.P. Mon. Not. R. Astron. Soc. 329, 775 (2002) GMRT observations of four suspected supernova remnants near the Galactic Centre.

J/MNRAS/329/877/ notes.dat table2.dat (43Kb) Caccianiga A., Marcha M.J., Anton S., Mack K.-H., Neeser M.J. Mon. Not. R. Astron. Soc. 329, 877 (2002) The CLASS blazar survey. - II. Optical properties.

J/MNRAS/330/1/  table3.dat  (11Kb)

Smith G.P., Smail I., Kneib J.-P., Czoske O., Ebeling H., Edge A.C., Pello R., Ivison R.J., Packham C., Le Borgne J.-F. Mon. Not. R. Astron. Soc. 330, 1 (2002) A Hubble Space Telescope lensing survey of X-ray-luminous galaxy clusters - II. A search for gravitationally lensed EROs.

J/MNRAS/330/75/  table1.dat table2.dat  (20Kb)

Contini T., Treyer M.A., Sullivan M., Ellis R.S. Mon. Not. R. Astron. Soc. 330, 75 (2002) Chemical abundances in a UV-selected sample of galaxies.

J/MNRAS/330/137/ table1.dat (17Kb) Noda S., Takeuti M., Abe F., Bond I.A., Dodd R.J., Hearnshaw J.B., Honda M., Honma M., Jugaku J., Kabe S., Kan-Ya Y., Kato Y., Kilmartin P.M., Matsubara Y., Masuda K., Muraki Y., Nakamura T., Nankivell G.R., Noguchi C., Ohnishi K., Reid M., Rattenbury N.J., Saito T., Sato H., Sekiguchi M., Skuljan J., Sullivan D.J., Sumi T., Watase Y., Wilkinson S., Yamada R., Yanagisawa T., Yock P.C.M., Yoshizawa M. Mon. Not. R. Astron. Soc. 330, 137 (2002) Study of variable stars in the MOA data base: long-period red variables in the Large Magellanic Cloud.

J/MNRAS/330/547/ table1.dat table2.dat table4.dat (30Kb) Terlevich A.I., Forbes D.A. Mon. Not. R. Astron. Soc. 330, 547 (2002) A catalogue and analysis of local galaxy ages and metallicities.

J/MNRAS/330/792/ table1.dat table2.dat (17Kb)

Kleyna J., Wilkinson M.I., Evans N.W., Gilmore G., Frayn C. Mon. Not. R. Astron. Soc. 330, 792 (2002) Dark matter in dwarf spheroidals - II. Observations and modelling of Draco.

J/MNRAS/331/45/  table5.dat table5.ps table5.tex  (543Kb)

Koen C., Eyer L. Mon. Not. R. Astron. Soc. 331, 45 (2002) New periodic variables from the Hipparcos epoch photometry.

J/MNRAS/331/474/ table1.dat table1b.dat (27Kb) Franco G.A.P. Mon. Not. R. Astron. Soc. 331, 474 (2002) On the infrared void in the Lupus dark clouds.

J/MNRAS/331/609/ fig/* moa.dat (6.2 Mb) Bayne G., Tobin W., Pritchard J.D., Bond I., Pollard K.R., Besier S.C., Noda S., Sumi T., Yanagisawa T., Sekiguchi M., Honda M., Muraki Y., Takeuti M., Hearnshaw J.B., Kilmartin P.M., Dodd R.J., Sullivan D.J., Yock P.C.M. Mon. Not. R. Astron. Soc. 331, 609 (2002) The MOA catalogue of eclipsing binary stars in the Small Magellanic Cloud.

J/MNRAS/332/422/ table1.dat (33Kb) Pierini D., Gavazzi G., Franzetti P., Scodeggio M., Boselli A. Mon. Not. R. Astron. Soc. 332, 422 (2002) 1.65-μm (H-band) surface photometry of galaxies. VIII. The near-IR κspace at z=0.

J/MNRAS/332/441/ table1.dat table2.dat table4.dat table5.dat (42Kb) Formiggini L., Brosch N., Almoznino E., Bowyer S., Lampton M. Mon. Not. R. Astron. Soc. 332, 441 (2002) Hidden subluminous stars among the FAUST UV sources towards Ophiuchus.

J/MNRAS/332/617 See VII/228/

J/MNRAS/333/100/ table23.dat (64Kb) Magliocchetti M., Maddox S.J., Jackson C.A., Bland-Hawthorn J., Bridges T., Cannon R., Cole S., Colless M., Collins C., Couch W., Dalton G., de Propris R., Driver S.P., Efstathiou G., Ellis R.S., Frenk C.S., Glazebrook K., Lahav O., Lewis I., Lumsden S., Peacock J.A., Peterson B.A., Sutherland W., Taylor K. Mon. Not. R. Astron. Soc. 333, 100 (2002) The 2dF Galaxy Redshift Survey: the population of nearby radio galaxies at the 1-mJy level.

J/MNRAS/333/423/ table2.dat (49Kb) Trentham N., Hodgkin S. Mon. Not. R. Astron. Soc. 333, 423 (2002) The luminosity function of the Virgo Cluster from M_{B} = -22 to -11.

J/MNRAS/334/20/  table1.dat table2.dat table7.dat  (63Kb)

Koen C., Kilkenny D., van Wyk F., Cooper D., Marang F. Mon. Not. R. Astron. Soc. 334, 20 (2002) UBV(RI)_{c} photometry of Hipparcos red stars.

J/MNRAS/335/10/  table1.dat  (46Kb)

Balogh M.L., Couch W.J., Smail I., Bower R.G., Glazebrook K. Mon. Not. R. Astron. Soc. 335, 10 (2002) An Hαsurvey of the rich cluster A 1689.

J/MNRAS/335/233/ clusters.dat refs.dat spectra/* template.dat (2.2 Mb) Piatti A.E., Bica E., Claria J.J., Santos J.F.C.Jr, Ahumada A.V. Mon. Not. R. Astron. Soc. 335, 233 (2002) Integrated spectral evolution of Galactic open clusters.

J/MNRAS/335/275/ table1.dat table2.dat (31Kb) Morris D.J., Hobbs G., Lyne A.G., Stairs I.H., Camilo F., Manchester R.N., Possenti A., Bell J.F., Kaspi V.M., D’Amico N., McKay N.P.F., Crawford F., Kramer M. Mon. Not. R. Astron. Soc. 335, 275 (2002) The Parkes Multi-Beam Pulsar Survey - II. Discovery and timing of 120 pulsars.

J/MNRAS/335/291/ table2.dat table3.dat table4.dat table5.dat table6.dat table7.dat (8.4 Mb) Naylor T., Totten E.J., Jeffries R.D., Pozzo M., Devey C.R., Thompson S.A. Mon. Not. R. Astron. Soc. 335, 291 (2002) Optimal photometry for colour-magnitude diagrams and its application to NGC 2547

J/MNRAS/335/555/ q0055.dat q0302.dat q1101.dat q1122.dat q1347.dat table1.dat (146Kb)

Kim T.-S., Carswell R.F., Cristiani S., D’odorico S., Giallongo E. Mon. Not. R. Astron. Soc., 335, 555-573 (2002) The physical properties of the Lyαforest at z 1.5 .

J/MNRAS/335/673/ tablea1.dat (17Kb) Vennes S., Smith R.J., Boyle B.J., Croom S.M., Kawka A., Shanks T., Miller L., Loaring N. Mon. Not. R. Astron. Soc. 335, 673 (2002) White dwarfs in the 2dF QSO Redshift Survey - I. Hydrogen-rich (DA) stars.

J/MNRAS/335/712/ tables.dat (32Kb) Trentham N., Tully R.B. Mon. Not. R. Astron. Soc. 335, 712 (2002) The faint end of the galaxy luminosity function.

J/MNRAS/335/1085/ table3.dat table4.dat table5.dat table6.dat (87Kb) Filipovic M.D., Bohlsen T., Reid W., Staveley-Smith L., Jones P.A., Nohejl K., Goldstein G. Mon. Not. R. Astron. Soc. 335, 1085 (2002) ATCA SMC Radio Continuum Source Catalogue. I. Source catalogues at 1.42, 2.37, 4.80 and 8.64 GHz.

J/MNRAS/336/879/ table3.dat (56Kb) Kotoneva E., Flynn C., Chiappini C., Matteucci F. Mon. Not. R. Astron. Soc. 336, 879 (2002) K dwarfs and the chemical evolution of the solar cylinder.

J/MNRAS/337/133/ be11.dat be11f.dat tr1.dat tr1f.dat (178Kb) Yadav R.K.S., Sagar R. Mon. Not. R. Astron. Soc. 337, 133 (2002) A deep UBVRI CCD photometric study of the open clusters Tr 1 and Be 11.

J/MNRAS/337/256/ table1.dat table4.dat (26Kb) Balogh M., Bower R.G., Smail I., Ziegler B.L., Davies R.L., Gaztelu A., Fritz A. Mon. Not. R. Astron. Soc. 337, 256 (2002) Galaxy properties in low X-ray luminosity clusters at z=0.25.

J/MNRAS/337/519/ table1.dat (13Kb) Glass I.S., Schultheis M. Mon. Not. R. Astron. Soc. 337, 519 (2002) M giants in MACHO, DENIS and ISOGAL.

J/MNRAS/337/641/ table1.dat (20Kb) Waugh M., Drinkwater M.J., Webster R.L., Staveley-Smith L., Kilborn V.A., Barnes D.G., Bhathal R., de Blok W.J.G., Boyce P.J., Disney M.J., Ekers R.D., Freeman K.C., Gibson B.K., Henning P.A., Jerjen H., Knezek P.M., Koribalski B., Marquarding M., Minchin R.F., Price R.M., Putman M.E., Ryder S.D., Sadler E.M., Stootman F., Zwaan M.A. Mon. Not. R. Astron. Soc. 337, 641 (2002) The large-scale distribution of neutral hydrogen in the Fornax region.

J/MNRAS/338/253/ extmap.dat (996Kb) Dutra C.M., Santiago B.X., Bica E.L.D., Barbuy B. Mon. Not. R. Astron. Soc. 338, 253 (2003) Extinction within 10 degrees of the Galactic Centre using 2MASS

J/MNRAS/338/508/ table2.dat (28Kb) Perez-Gonzalez P.G., Gil de Paz A., Zamorano J., Gallego J., Alonso-Herrero A., Aragon-Salamanca A. Mon. Not. R. Astron. Soc. 338, 508 (2003) Stellar populations in local star-forming galaxies. I. Data and modelling procedure.

J/MNRAS/338/525/ table2.dat (33Kb) Perez-Gonzalez P.G., Gil De Paz A., Zamorano J., Gallego J., Alonso-Herrero A., Aragon-Salamanca A. Mon. Not. R. Astron. Soc., 338, 525-543 (2003) Stellar populations in local star-forming galaxies - II. Recent star formation properties and stellar masses.

J/MNRAS/339/105/ table1.dat (16Kb) Muller E., Staveley-Smith L., Zealey W., Stanimirovic S. Mon. Not. R. Astron. Soc. 339, 105 (2003) High-resolution HI observations of the Western Magellanic Bridge.

J/MNRAS/339/652/ catalog.dat (359Kb) Kaldare R., Colless M., Raychaudhury S., Peterson B.A. Mon. Not. R. Astron. Soc. 339, 652 (2003) FLASH redshift survey - I. Observations and catalogue.

J/MNRAS/339/701/ vhlmc.dat (68Kb) Zhao H., Ibata R.A., Lewis G.F., Irwin M.J. Mon. Not. R. Astron. Soc. 339, 701 (2003) Kinematic outliers in the Large Magellanic Cloud: constraints on star-star microlensing.

J/MNRAS/339/913/ refs.dat table2.dat (20Kb) Edge A.C., Ebeling H., Bremer M., Rottgering H., Van Haarlem M.P., Rengelink R., Courtney N.J.D. Mon. Not. R. Astron. Soc. 339, 913 (2003) The discovery of two distant, massive clusters of galaxies in the ROSAT All-Sky Survey.

J/MNRAS/340/304/ table1.dat table5.dat table6.dat (53Kb) Reddy B.E., Tomkin J., Lambert D.L., Prieto C.A. Mon. Not. R. Astron. Soc. 340, 304 (2003) The chemical compositions of Galactic disc F and G dwarfs.

J/MNRAS/340/1249/ table1.dat table2.dat table3.dat (139Kb) Piatti A.E., Claria J.J., Ahumada A.V. Mon. Not. R. Astron. Soc. 340, 1249 (2003) The relatively young and metal-poor Galactic open cluster NGC 2194.

J/MNRAS/340/1346/ bulsc1.dat (4.8 Mb) Sumi T., Eyer L., Wozniak P.R. Mon. Not. R. Astron. Soc. 340, 1346 (2003) Measurements of streaming motions of the Galactic bar with red clump giants.

J/MNRAS/341/1 See VIII/72/

J/MNRAS/341/13 See VIII/72/

J/MNRAS/341/534/ table1.dat (14Kb) Morgan D.H., Cannon R.D., Hatzidimitriou D., Croke B.F.W. Mon. Not. R. Astron. Soc. 341, 534 (2003) J-type carbon stars in the Large Magellanic Cloud.

J/MNRAS/341/805/ catalog.dat flags.dat (1.6 Mb) Pozzo M., Naylor T., Jeffries R.D., Drew J.E. Mon. Not. R. Astron. Soc., 341, 805-822 (2003) The discovery of low-mass pre-main-sequence stars in Cepheus OB3b.

J/MNRAS/341/993/ table1.dat (14Kb) Cirasuolo M., Magliocchetti M., Celotti A., Danese L. Mon. Not. R. Astron. Soc. 341, 993 (2003) The radio-loud/radio-quiet dichotomy: news from the 2dF QSO Redshift Survey.

J/MNRAS/341/1066/ table1.dat table3.dat (9Kb) Taylor A.C., Carreira P., Cleary K., Davies R.D., Davis R.J., Dickinson C., Grainge K., Gutierrez C.M., Hobson M.P., Jones M.E., Kneissl R., Lasenby A., Leahy J.P., Maisinger K., Pooley G.G., Rebolo R., Rubinomartin J.A., Rusholme B., Saunders R.D.E., Savage R., Scott P.F., Slosar A., Sosa Molina P.J., Titterington D., Waldram E., Watson R.A., Wilkinson A. Mon. Not. R. Astron. Soc., 341, 1066-1075 (2003) First results from the Very Small Array - II. Observations of the cosmic microwave background.

J/MNRAS/341/1093/ clusters.dat extsrcs.dat fields.dat numflux.dat selfn.dat (59Kb) Burke D.J., Collins C.A., Sharples R.M., Romer A.K., Nichol R.C. Mon. Not. R. Astron. Soc. 341, 1093 (2003) The Southern SHARC catalogue: a ROSAT survey for distant galaxy clusters.

J/MNRAS/342/575/ table1.dat (33Kb) Ciliegi P., Vignali C., Comastri A., Fiore F., La Franca F., Perola G.C. Mon. Not. R. Astron. Soc. 342, 575 (2003) The BeppoSAX High Energy Large Area Survey (HELLAS). VI. The radio properties.

J/MNRAS/342/738/ table2.dat (29Kb) Lang R.H., Boyce P.J., Kilborn V.A., Minchin R.F., Disney M.J., Jordan C.A., Grossi M., Garcia D.A., Freeman K.C., Phillipps S., Wright A.E. Mon. Not. R. Astron. Soc. 342, 738 (2003) First results from the HI Jodrell All Sky Survey: inclination-dependent selection effects in a 21-cm blind survey.

J/MNRAS/342/802/ notes.dat table2.dat table5.dat tablea1.dat (46Kb) McHardy I.M., Gunn K.F., Newsam A.M., Mason K.O., Page M.J., Takata T., Sekiguchi K., Sasseen T., Cordova F., Jones L.R., Loaring N.

Mon. Not. R. Astron. Soc. 342, 802 (2003) A medium-deep Chandra and Subaru survey of the 13-h XMM/ROSAT deep survey area.

J/MNRAS/342/915/ table4.dat table5.dat table6.dat (24Kb) Waldram E.M., Pooley G.G., Grainge K.J.B., Jones M.E., Saunders R.D.E., Scott P.F., Taylor A.C. Mon. Not. R. Astron. Soc. 342, 915 (2003) 9C: a survey of radio sources at 15 GHz with the Ryle Telescope.

J/MNRAS/342/1117 See VIII/70A/

J/MNRAS/342/1241/ tablea1.dat tablea2.dat tablea3.dat tablea4.dat (35Kb) Pinfield D.J., Dobbie P.D., Jameson R.F., Steele I.A., Jones H.R.A., Katsiyannis A.C. Mon. Not. R. Astron. Soc. 342, 1241 (2003) Brown dwarfs and low-mass stars in the Pleiades and Praesepe: membership and binarity.

J/MNRAS/342/1299/ refs.dat table1.dat table2.dat table3.dat table4.dat table5.dat (75Kb) Kramer M., Bell J.F., Manchester R.N., Lyne A.G., Camilo F., Stairs I.H., D’Amico N., Kaspi V.M., Hobbs G., Morris D.J., Crawford F., Possenti A., Joshi B.C., McLaughlin M.A., Lorimer D.R., Faulkner A.J. Mon. Not. R. Astron. Soc. 342, 1299 (2003) The Parkes Multibeam Pulsar Survey - III. Young pulsars and the discovery and timing of 200 pulsars.

J/MNRAS/343/293/ list.dat table1.dat table2.dat (252Kb) Manners J.C., Johnson O., Almaini O., Willott C.J., Gonzalez-Solares E., Lawrence A., Mann R.G., Perez-Fournon I., Dunlop J.S., Mcmahon R.G., Oliver S.J., Rowan-Robinson M., Serjeant S. Mon. Not. R. Astron. Soc., 343, 293-305 (2003) The ELAIS deep X-ray survey. I. Chandra source catalogue and first results.

J/MNRAS/343/665/ refs.dat table1.dat tableb1.dat tableb2.dat tableb3.dat tableb4.dat tableb5.dat (27Kb) Goudfrooij P., Strader J., Brenneman L., Kissler-Patig M., Minniti D., Huizinga J.E. Mon. Not. R. Astron. Soc. 343, 665 (2003) Hubble Space Telescope observations of globular cluster systems along the Hubble sequence of spiral galaxies.

J/MNRAS/343/1131/ position.dat table1.dat table2.dat table3.dat (30Kb) Tsapras Y., Horne K., Kane S., Carson R. Mon. Not. R. Astron. Soc. 343, 1131 (2003) Microlensing limits on numbers and orbits of extrasolar planets from the 1998-2000 OGLE events.

J/MNRAS/344/134/ table3.dat table4.dat table5.dat table12.dat (22Kb) Humphrey P.J., Fabbiano G., Elvis M., Church M.J., Balucinska-Church M. Mon. Not. R. Astron. Soc. 344, 134 (2003) Ultraluminous X-ray source populations in normal galaxies: a preliminary survey with Chandra.

J/MNRAS/344/181/ table1.dat table2.dat (33Kb) Coziol R. Mon. Not. R. Astron. Soc. 344, 181 (2003) The PDS versus Markarian starburst galaxies: comparing strong and weak IRAS emitter at 12 and 25 μm in the nearby Universe.

J/MNRAS/344/188/ table3.dat table4.dat (23Kb) Karick A.M., Drinkwater M.J., Gregg M.D. Mon. Not. R. Astron. Soc. 344, 188 (2003) The surface brightness and colour-magnitude relations for Fornax cluster galaxies.

J/MNRAS/344/283/ table2.dat (17Kb) Brand K., Rawlings S., Hill G.J., Lacy M., Mitchell E., Tufts J. Mon. Not. R. Astron. Soc. 344, 283 (2003) Two 100-Mpc-scale structures in the three-dimensional distribution of radio galaxies and their implications.

J/MNRAS/344/307 See VII/240/

J/MNRAS/344/527 See VI/112/

J/MNRAS/344/602/ table1.dat (9Kb) Basilakos S. Mon. Not. R. Astron. Soc. 344, 602 (2003) Shape statistics of Sloan Digital Sky Survey superclusters.

J/MNRAS/344/887/ catalog.dat (184Kb) Serjeant S., Dunlop J.S., Mann R.G., Rowan-Robinson M., Hughes D., Efstathiou A., Blain A., Fox M., Ivison R.J., Jenness T., Lawrence A., Longair M., Oliver S., Peacock J.A. Mon. Not. R. Astron. Soc., 344, 887-904 (2003) Submillimetre observations of the Hubble Deep Field and Flanking Fields.

J/MNRAS/345/1/  tablea1.dat  (11Kb)

Moller O., Hewett P., Blain A.W. Mon. Not. R. Astron. Soc. 345, 1 (2003) Discs in early-type lensing galaxies: effects on magnification ratios and measurements of H_{0}.

J/MNRAS/345/609/ table2.dat table3.dat (21Kb) Murphy M.T., Webb J.K., Flambaum V.V. Mon. Not. R. Astron. Soc., 345, 609-638 (2003) Further evidence for a variable fine-structure constant from Keck/HIRES QSO absorption spectra.

J/MNRAS/345/747/ table2.dat table3.dat (32Kb) Mackey A.D., Gilmore G.F. Mon. Not. R. Astron. Soc., 345, 747-761 (2003) RR Lyrae stars in four globular clusters in the Fornax dwarf galaxy.

J/MNRAS/345/949/ tablea1.dat (13Kb) Whitlock S., Forbes D.A., Beasley M.A. Mon. Not. R. Astron. Soc., 345, 949-959 (2003) UBRI photometry of globular clusters in the Leo group galaxy NGC 3379.

J/MNRAS/346/18/  saurer1.dat saurer2.dat saurer3.dat  (340Kb)

Carraro G., Baume G. Mon. Not. R. Astron. Soc., 346, 18-26 (2003) The old open clusters Saurer A, B and C revisited.

J/MNRAS/346/390/ table1.dat table2.dat table3.dat (242Kb) Piatti A.E., Claria J.J., Ahumada A.V. Mon. Not. R. Astron. Soc., 346, 390-402 (2003) Intermediate-age Galactic open clusters: fundamental parameters of NGC 2627.

J/MNRAS/346/627/ notes.dat tablea1.dat tablea2.dat tableb1.dat tablec1.dat (79Kb) Best P.N., Arts J.N., Rottgering H.J.A., Rengelink R., Brookes M.H., Wall J. Mon. Not. R. Astron. Soc., 346, 627-683 (2003) CENSORS: A Combined EIS-NVSS Survey Of Radio Sources. I. Sample definition, radio data and optical identifications.

J/MNRAS/346/787/ table4.dat table4.ps (82Kb) Minchin R.F., Disney M.J., Boyce P.J., De Blok W.J.G., Parker Q.A., Banks G.D., Freeman K.C., Garcia D.A., Gibson B.K., Grossi M., Haynes R.F., Knezek P.M., Lang R.H., Malin D.F., Price R.M., Stewart I.M., Wright A.E. Mon. Not. R. Astron. Soc., 346, 787-802 (2003) HIDEEP - an extragalactic blind survey for very low column-density neutral hydrogen.

J/MNRAS/346/1103/ refs.dat table1.dat (50Kb) Peroux C., Mcmahon R.G., Storrie-lombardi L.J., Irwin M.J. Mon. Not. R. Astron. Soc., 346, 1103-1115 (2003) The evolution of Ω(HI) and the epoch of formation of damped Lymanαabsorbers.

J/MNRAS/346/1125/ table4.dat (23Kb) Gilbank D.G., Smail I., Ivison R.J., Packham C. Mon. Not. R. Astron. Soc., 346, 1125-1142 (2003) An HST morphological survey of faint extremely red objects.

J/MNRAS/346/1143/ table1.dat table2.dat table3.dat (2.8 Mb) Burningham B., Naylor T., Jeffries R.D., Devey C.R. Mon. Not. R. Astron. Soc., 346, 1143-1150 (2003) On the nature of Collinder 121: insights from the low-mass pre-main sequence.

J/MNRAS/347/837/ fits/* notes.txt table1.dat table2.dat (761.1 Mb) De Breuck C., Hunstead R.W., Sadler E.M., Rocca-Volmerange B., Klamer I. Mon. Not. R. Astron. Soc., 347, 837-853 (2004) A search for distant radio galaxies from SUMSS and NVSS - I. Sample definition, radio and K-band imaging.

J/MNRAS/348/857/ table1.dat table2.dat (16Kb) Holt J., Benn C.R., Vigotti M., Pedani M., Carballo R., Gonzalez-serrano J.I., Mack K.-H., Garcia B. Mon. Not. R. Astron. Soc., 348, 857-865 (2004) A sample of radio-loud QSOs at redshift 4.

J/MNRAS/348/937/ table1.dat (21Kb) Caccianiga A., Marcha M.J.M. Mon. Not. R. Astron. Soc., 348, 937-954 (2004)

The CLASS blazar survey: testing the blazar sequence.

J/MNRAS/348/1439/ pm/* table1.dat table3.dat (233.6 Mb) Sumi T., Wu X., Udalski A., Szymanski M., Kubiak M., Pietrzynski G., Soszynski I., Wozniak P., Zebrun K., Szewczyk O., Wyrzykowski L. Mon. Not. R. Astron. Soc., 348, 1439-1450 (2004) The Optical Gravitational Lensing Experiment: catalogue of stellar proper motions in the OGLE-II Galactic bulge fields.

J/MNRAS/349/135/ table1.dat (8Kb) Georgakakis A., Georgantopoulos I., Vallbe M., Kolokotronis V., Basilakos S., Plionis M., Stewart G.C., Shanks T., Boyle B.J. Mon. Not. R. Astron. Soc., 349, 135-145 (2004) The XMM-Newton/2dF survey - II. The nature of X-ray-faint optically bright X-ray sources.

J/MNRAS/349/641/ table1.dat table2.dat (424Kb) Piatti A.E., Claria J.J., Ahumada A.V. Mon. Not. R. Astron. Soc., 349, 641-648 (2004) The old open cluster Trumpler 5: a reddened, metal-poor anticentre cluster.

J/MNRAS/349/757/ table1.dat (28Kb) Lambert D.L., Reddy B.E. Mon. Not. R. Astron. Soc., 349, 757-767 (2004) Lithium abundances of the local thin disc stars.

J/MNRAS/349/793/ refs.dat sample.dat table2.dat table3.dat (101Kb) Perinotto M., Morbidelli L., Scatarzi A. Mon. Not. R. Astron. Soc., 349, 793-815 (2004) A reanalysis of chemical abundances in galactic PNe and comparison with theoretical predictions.

J/MNRAS/349/833/ table1.dat table2.dat table3.dat (66Kb) Bayne G.P., Tobin W., Pritchard J.D., Pollard K.R., Albrow M.D. Mon. Not. R. Astron. Soc., 349, 833-840 (2004) CCD photometry of variable stars in the Magellanic Clouds. VII. The eclipsing binaries MACHO*05:36:48.7-69:17:00 in the LMC, and MOA J005018.4-723855 and J005623.5-722123 in the SMC.

J/MNRAS/349/1059/ atype.dat btype.dat field.dat (2.9 Mb) Wray J.J., Eyer L., Paczynski B. Mon. Not. R. Astron. Soc., 349, 1059-1068 (2004) OGLE small-amplitude variables in the Galactic bar.

J/MNRAS/349/1069/ refs.dat table1.dat (45Kb) Karatas Y., Bilir S., Eker Z., Demircan O. Mon. Not. R. Astron. Soc., 349, 1069-1092 (2004) Kinematics of chromospherically active binaries and evidence of an orbital period decrease in binary evolution.

J/MNRAS/349/1278/ chip1.dat chip2.dat chip3.dat chip6.dat chip7.dat chip8.dat (14.2 Mb) Monaco L., Pancino E., Ferraro F.R., Bellazzini M. Mon. Not. R. Astron. Soc., 349, 1278-1290 (2004) Wide-field photometry of the Galactic globular cluster M22.

J/MNRAS/349/1397 See VII/241/

J/MNRAS/350/679/ table1.dat (13Kb) Mennickent R.E., Pietrzynski G., Gieren W. Mon. Not. R. Astron. Soc., 350, 679-684 (2004) The Araucaria Project. Variable stars outside the Local Group: NGC 300.

J/MNRAS/350/785/ tablea12.dat tablea34.dat tablea56.dat (73Kb) Waskett T.J., Eales S.A., Gear W.K., McCracken H.J., Brodwin M., Nandra K., Laird E.S., Lilly S. Mon. Not. R. Astron. Soc., 350, 785-797 (2004) XMM-Newton surveys of the Canada-France Redshift Survey fields - II. The X-ray catalogues, the properties of the host galaxies and the redshift distribution.

J/MNRAS/350/1195 See VIII/73/

J/MNRAS/350/1485/ table1.dat (23Kb) Magliocchetti M., Maddox S.J., Hawkins E., Peacock J.A., Bland-hawthorn J., Bridges T., Cannon R., Cole S., Colless M., Collins C., Couch W., Dalton G., De Propris R., Driver S.P., Efstathiou G., Ellis R.S., Frenk C.S., Glazebrook K., Jackson C.A., Jones B., Lahav O., Lewis I., Lumsden S., Norberg P., Peterson B.A., Sutherland W., Taylor K., (the 2DFGRS Team) Mon. Not. R. Astron. Soc., 350, 1485-1494 (2004)

The 2dF Galaxy Redshift Survey: clustering properties of radio galaxies.

J/MNRAS/351/31/  tablea1.dat  (64Kb)

Hands A.D.P., Warwick R.S., Watson M.G., Helfand D.J. Mon. Not. R. Astron. Soc., 351, 31-56 (2004) X-ray source populations in the Galactic plane.

J/MNRAS/351/83/  table1.dat table2.dat table3.dat table4.dat  (31Kb)

Landt H., Padovani P., Perlman E.S., Giommi P. Mon. Not. R. Astron. Soc., 351, 83-100 (2004) A physical classification scheme for blazars.

J/MNRAS/351/265/ table1.dat (9Kb) Desai V., Dalcanton J.J., Mayer L., Reed D., Quinn T., Governato F. Mon. Not. R. Astron. Soc., 351, 265-278 (2004) The cluster galaxy circular velocity function.

J/MNRAS/351/1071/ table23.dat (2.4 Mb) An J.H., Evans N.W., Hewett P., Baillon P., Novati S.C., Carr B.J., Creze M., Giraud-heraud Y., Gould A., Jetzer P., Kaplan J., Kerins E., Paulin-henriksson S., Smartt S.J., Stalin C.S., Tsapras Y. Mon. Not. R. Astron. Soc., 351, 1071-1098 (2004) The POINT-AGAPE Survey - I. The variable stars in M31.

J/MNRAS/351/1204/ m15.dat m30.dat ngc288.dat ngc362.dat ngc6342.dat ngc6380.dat ngc6440.dat ngc6441.dat ngc6624.dat ngc6752.dat (1.6 Mb) Valenti E., Ferraro F.R., Origlia L. Mon. Not. R. Astron. Soc. 351, 1204 (2004) Red giant branch in the near-infrared colour-magnitude diagrams. I: Calibration of photometric indices.

J/MNRAS/351/1290/ catalog.dat unassoc.dat (1.8 Mb) Rowan-Robinson M., Lari C., Perez-Fournon I., Gonzalez-Solares E.A., Franca F.L., Vaccari M., Oliver S., Gruppioni C., Ciliegi P., Heraudeau P., Serjeant S., Efstathiou A., Babbedge T., Matute I., Pozzi F., Franceschini A., Vaisanen P., Afonso-Luis A., Alexander D.M., Almaini O., Baker A.C., Basilakos S., Barden M., Del Burgo C., Bellas-Velidis I., Cabrera-Guerra F., Carballo R., Cesarsky C.J., Clements D.L., Crockett H., Danese L., Dapergolas A., Drolias B., Eaton N., Egami E., Elbaz D., Fadda D., Fox M., Genzel R., Goldschmidt P., Gonzalez-Serrano J.I., Graham M., Granato G.L., Hatziminaoglou E., Herbstmeier U., Joshi M., Kontizas E., Kontizas M., Kotilainen J.K., Kunze D., Lawrence A., Lemke D., Linden-Vornle M.J.D., Mann R.G., Marquez I., Masegosa J., Mcmahon R.G., Miley G., Missoulis V., Mobasher B., Morel T., Norgaard-Nielsen H., Omont A., Papadopoulos P., Puget J.-L., Rigopoulou D., Rocca-Volmerange B., Sedgwick N., Silva L., Sumner T., Surace C., Vila-Vilaro B., Van Der Werf P., Verma A., Vigroux L., Villar-Martin M., Willott C.J., Carraminana A., Mujica R. Mon. Not. R. Astron. Soc., 351, 1290-1306 (2004) The European Large-Area ISO Survey (ELAIS): the final band-merged catalogue.

J/MNRAS/351/1401/ tablea1.dat (12.0 Mb) Jeffries R.D., Naylor T., Devey C.R., Totten E.J. Mon. Not. R. Astron. Soc., 351, 1401-1422 (2004) Low mass stars, brown dwarf candidates and the mass function of the young open cluster NGC 2547.

J/MNRAS/352/91/  table2.dat  (11Kb)

Georgantopoulos I., Georgakakis A., Akylas A., Stewart G.C., Giannakis O., Shanks T., Kitsionas S. Mon. Not. R. Astron. Soc., 352, 91-100 (2004) The XMM-Newton/2dF survey - IV. The X-ray spectral properties of the hard sources.

J/MNRAS/352/131/ tablea1.dat (34Kb) Seymour N., Mchardy I.M., Gunn K.F. Mon. Not. R. Astron. Soc., 352, 131-141 (2004) Radio observations of the 13^{h}XMM-Newton/ROSAT Deep X-ray Survey Area.

J/MNRAS/352/285/ table4.dat table7.dat table11.dat (73Kb) Clewley L., Warren S.J., Hewett P.C., Norris J.E., Evans N.W. Mon. Not. R. Astron. Soc., 352, 285-298 (2004) Distant field blue horizontal branch stars and the mass of the Galaxy - II. Photometry and spectroscopy of UKST candidates 16 B 19.5, 11 R 52kpc

J/MNRAS/352/478/ table1.dat table2.dat table3.dat (17Kb) Roberts S., Davies J., Sabatini S., Van Driel W., O’neil K., Baes M., Linder S., Smith R., Evans R. Mon. Not. R. Astron. Soc., 352, 478-492 (2004) A search for low surface brightness dwarf galaxies in different environments.

J/MNRAS/352/939/ ngp.dat sgp.dat table1.dat table2.dat table3.dat voids.dat (1.7 Mb)

Erdogdu P., Lahav O., Zaroubi S., Efstathiou G., Moody S., Peacock J.A., Colless M., Baldry I.K., Baugh C.M., Bland-hawthorn J., Bridges T., Cannon R., Cole S., Collins C., Couch W., Dalton G., De Propris R., Driver S.P., Ellis R.S., Frenk C.S., Glazebrook K., Jackson C., Lewis I., Lumsden S., Maddox S., Madgwick D., Norberg P., Peterson B.A., Sutherland W., Taylor K. Mon. Not. R. Astron. Soc., 352, 939-960 (2004) The 2dF Galaxy Redshift Survey: Wiener reconstruction of the cosmic web.

J/MNRAS/352/1245/ table3.dat table4.dat (34Kb) Chan B.H.P., Cram L.E., Sadler E.M., Killeen N.E.B., Jackson C.A., Mobasher B., Ekers R.D. Mon. Not. R. Astron. Soc., 352, 1245-1254 (2004) Faint 1.4-GHz radio sources in the 2dF Galaxy Redshift Survey.

J/MNRAS/352/1439/ refs.dat table1.dat table2.dat table4.dat table5.dat table7.dat table8.dat table9.dat (103Kb) Hobbs G., Faulkner A., Stairs I.H., Camilo F., Manchester R.N., Lyne A.G., Kramer M., D’amico N., Kaspi V.M., Possenti A., Mclaughlin M.A., Lorimer D.R., Burgay M., Joshi B.C., Crawford F. Mon. Not. R. Astron. Soc., 352, 1439-1472 (2004) The Parkes Multi-Beam Pulsar Survey - IV. Discovery of 180 pulsars and parameters for 281 previously known pulsars.

J/MNRAS/353/211/ table4.dat (9Kb) Carballo R., Cofino A.S., Gonzalez-serrano J.I. Mon. Not. R. Astron. Soc., 353, 211-220 (2004) Selection of quasar candidates from combined radio and optical surveys using neural networks.

J/MNRAS/353/310/ table1.dat (83Kb) Rodriguez E., Garcia J.M., Gamarova A.Y., Costa V., Daszynska-Daszkiewicz J., Lopez-Gonzalez M.J., Mkrtichian D.E., Rolland A. Mon. Not. R. Astron. Soc. 353, 310 (2004) Delta Scuti-type pulsations in eclipsing binary systems: AB Cas.

J/MNRAS/353/589/ notes.dat table1.dat (25Kb) Phillips J.P. Mon. Not. R. Astron. Soc., 353, 589-600 (2004) Planetary nebula distances re-examined: an improved statistical scale.

J/MNRAS/353/601/ notes.dat tab2df.dat tabmatch.dat (1.1 Mb) Evans C.J., Howarth I.D., Irwin M.J., Burnley A.W., Harries T.J. Mon. Not. R. Astron. Soc. 353, 601 (2004) A 2dF survey of the Small Magellanic Cloud.

J/MNRAS/353/705/ lmcvar.dat smcvar.dat (892Kb) Ita Y., Tanabe T., Matsunaga N., Nakajima Y., Nagashima C., Nagayama T., Kato D., Kurita M., Nagata T., Sato S., Tamura M., Nakaya H., Nakada Y. Mon. Not. R. Astron. Soc., 353, 705-712 (2004) Variable stars in the Magellanic Clouds - II. The data and infrared properties.

J/MNRAS/353/1025/ table2.dat table3.dat table4.dat table5.dat (128Kb) Nanda Kumar M.S., Kamath U.S., Davis C.J. Mon. Not. R. Astron. Soc., 353, 1025-1034 (2004) Embedded star clusters in the W51 giant molecular cloud.

J/MNRAS/353/1311/ table1.dat table2.dat table4.dat table6.dat (147Kb) Hobbs G., Lyne A.G., Kramer M., Martin C.E., Jordan C. Mon. Not. R. Astron. Soc., 353, 1311-1344 (2004) Long-term timing observations of 374 pulsars.

J/MNRAS/354/L25/ table1.dat (22Kb) Bouche N., Murphy M.T., Peroux C. Mon. Not. R. Astron. Soc., 354, L25-L29 (2004) The clustering of luminous red galaxies around MgII absorbers.

J/MNRAS/354/L31/ table1.dat (195Kb) Murphy M.T., Liske J. Mon. Not. R. Astron. Soc., 354, L31-L36 (2004) Dust-reddening and gravitational lensing of SDSS QSOs due to foreground damped Lyman αsystems.

J/MNRAS/354/305/ table2.dat table3.dat table5.dat (46Kb) Ricci R., Sadler E.M., Ekers R.D., Staveley-smith L., Wilson W.E., Kesteven M.J., Subrahmanyan R., Walker M.A., Jackson C.A., De Zotti G. Mon. Not. R. Astron. Soc., 354, 305-320 (2004)

First results from the Australia Telescope Compact Array 18-GHz pilot survey.

J/MNRAS/354/485/ table2.dat (33Kb) Bolton R.C., Cotter G., Pooley G.G., Riley J.M., Waldram E.M., Chandler C.J., Mason B.S., Pearson T.J., Readhead A.C.S. Mon. Not. R. Astron. Soc., 354, 485-521 (2004) The radio source population at high frequency: follow-up of the 15-GHz 9C survey

J/MNRAS/354/708/ table.dat (1.1 Mb) Bellazzini M., Gennari N., Ferraro F.R., Sollima A. Mon. Not. R. Astron. Soc., 354, 708-712 (2004) The distance to the Leo I dwarf spheroidal galaxy from the red giant branch tip.

J/MNRAS/354/980/ table1.dat (9Kb) Nagayama T., Woudt P.A., Nagashima C., Nakajima Y., Kato D., Kurita M., Nagata T., Nakaya H., Tamura M., Sugitani K., Wakamatsu K., Sato S. Mon. Not. R. Astron. Soc., 354, 980-990 (2004) A deep near-infrared survey around the giant radio galaxy PKS 1343-601.

J/MNRAS/355/20/  notes.dat table1.dat table2.dat table3.dat  (24Kb)

Jarvis M.J., Cruz M.J., Cohen A.S., Rottgering H.J.A., Kassim N.E. Mon. Not. R. Astron. Soc., 355, 20-30 (2004) Near-infrared K-band imaging of a sample of ultra-steep-spectrum radio sources selected at 74 MHz.

J/MNRAS/355/44/  table2.dat  (191Kb)

Payne J.L., Filipovic M.D., Reid W., Jones P.A., Staveley-Smith L., White G.L. Mon. Not. R. Astron. Soc. 355, 44 (2004) An ATCA radio-continuum study of the Small Magellanic Cloud. II. Source identification and classification.

J/MNRAS/355/82/  table2.dat  (28Kb)

Trujillo I., Aguerri J.A.L. Mon. Not. R. Astron. Soc., 355, 82-96 (2004) Quantitative morphological analysis of the Hubble Deep Field North and Hubble Deep Field South - I. Earlyand late-type luminosity-size relations of galaxies out to z 1.

J/MNRAS/355/97/  table1.dat table2.dat  (14Kb)

Manners J.C., Serjeant S., Bottinelli S., Vaccari M., Franceschini A., Perez-Fournon I., Gonzalez-Solares E., Willott C.J., Johnson O., Almaini O., Rowan-Robinson M., Oliver S. Mon. Not. R. Astron. Soc., 355, 97-105 (2004) Mid-infrared sources in the ELAIS Deep X-ray Survey.

J/MNRAS/355/307/ sa114.dat (836Kb) Karaali S., Bilir S., Hamzaoglu E. Mon. Not. R. Astron. Soc. 355, 307 (2004) A different approach for the estimation of Galactic model parameters.

J/MNRAS/355/585/ notes.dat table1.dat (127Kb) Greaves J. Mon. Not. R. Astron. Soc., 355, 585-590 (2004) New Northern hemisphere common proper-motion pairs.

J/MNRAS/355/608/ table1.dat tablea1.dat (173Kb) Forbes D.A., Faifer F.R., Forte J.C., Bridges T., Beasley M.A., Gebhardt K., Hanes D.A., Sharples R., Zepf S.E. Mon. Not. R. Astron. Soc., 355, 608-616 (2004) Gemini/GMOS imaging of globular clusters in the Virgo galaxy NGC 4649 (M60).

J/MNRAS/355/959/ table3.dat (17Kb) Dodd R.J. Mon. Not. R. Astron. Soc., 355, 959-972 (2004) Data mining in the young open cluster IC 2391.

J/MNRAS/355/1196/ table1.dat table2.dat (14Kb) Morgan D.H., Hatzidimitriou D., Cannon R.D. Mon. Not. R. Astron. Soc., 355, 1196-1206 (2004) Merrill-Sanford bands in Large Magellanic Cloud carbon stars.

J/MNRAS/355/1237/ table1.dat (7Kb) Hagele G.F., Albacete Colombo J.F., Barba R.H., Bosch G.L. Mon. Not. R. Astron. Soc., 355, 1237-1243 (2004) G287.84-0.82: an infrared star cluster in the Carina nebula.

J/MNRAS/355/1251/ table2.dat (15Kb) Laurikainen E., Salo H., Buta R., Vasylyev S. Mon. Not. R. Astron. Soc., 355, 1251-1271 (2004) Bar-induced perturbation strengths of the galaxies in the Ohio State University Bright Galaxy Survey - I.

J/MNRAS/355/1272/ table1.dat table2.dat (22Kb) Maheswar G., Manoj P., Bhatt H.C. Mon. Not. R. Astron. Soc., 355, 1272-1282 (2004) Photometric distances to dark clouds: cometary globule CG 12.

J/MNRAS/355/1303/ table1.dat table2.dat (20Kb) Minchin R.F., Disney M.J., Parker Q.A., Boyce P.J., De Blok W.J.G., Banks G.D., Ekers R.D., Freeman K.C., Garcia D.A., Gibson B.K., Grossi M., Haynes R.F., Knezek P.M., Lang R.H., Malin D.F., Price R.M., Putman M., Stewart I.M., Wright A.E. Mon. Not. R. Astron. Soc., 355, 1303-1314 (2004) The cosmological significance of low surface brightness galaxies found in a deep blind neutral hydrogen survey.

J/MNRAS/356/89/  catalog.dat flags.dat members.dat nonmemb.dat uncert.dat  (3.5 Mb)

Kenyon M.J., Jeffries R.D., Naylor T., Oliveira J.M., Maxted P.F.L. Mon. Not. R. Astron. Soc., 356, 89-106 (2005) Membership, binarity and accretion among very low-mass stars and brown dwarfs of the σOrionis cluster.

J/MNRAS/356/331/ table2.dat (14Kb) Sumi T., Wozniak P.R., Eyer L., Dobrzycki A., Udalski A., Szymanski M.K., Kubiak M., Pietrzynski G., Soszynski I., Zebrun K., Szewczyk O., Wyrzykowski L. Mon. Not. R. Astron. Soc., 356, 331-335 (2005) Variability-selected QSO candidates in OGLE-II Galactic bulge fields.

J/MNRAS/356/545/ table1.dat table3.dat table4a.dat table4b.dat (36Kb) Fremat Y., Lampens P., Hensberge H. Mon. Not. R. Astron. Soc., 356, 545-556 (2005) Spectral disentangling of the triple system DG Leo: orbits and chemical composition.

J/MNRAS/356/568/ table3.dat tablea1.dat tablea2.dat (56Kb) Nandra K., Laird E.S., Adelberger K., Gardner J.P., Mushotzky R.F., Rhodes J., Steidel C.C., Teplitz H.I., Arnaud K.A. Mon. Not. R. Astron. Soc., 356, 568-586 (2005) A deep Chandra survey of the Groth Strip - I. The X-ray data.

J/MNRAS/356/1256/ table6.dat (36Kb) Caliskan H. Mon. Not. R. Astron. Soc. 356, 1256 (2005) Elemental abundance analysis of magnetic chemically peculiar star HR 8216.

J/MNRAS/356/1583/ table3.dat (12Kb) Burningham B., Naylor T., Littlefair S.P., Jeffries R.D. Mon. Not. R. Astron. Soc., 356, 1583-1591 (2005) Contamination and exclusion in the σOrionis young group.

J/MNRAS/357/265/ catalog.dat (7.8 Mb) Sollima A., Ferraro F.R., Pancino E., Bellazzini M. Mon. Not. R. Astron. Soc. 357, 265 (2005) On the discrete nature of the Red Giant Branch of omega Centauri.

J/MNRAS/357/345/ tablea1.dat (21Kb) Zwintz K., Marconi M., Reegen P., Weiss W.W. Mon. Not. R. Astron. Soc., 357, 345-353 (2005) Search for pulsating pre-main-sequence stars in NGC 6383.

J/MNRAS/357/401/ table1.dat table2.dat (30Kb) Jenkins L.P., Roberts T.P., Warwick R.S., Kilgard R.E., Ward M.J. Mon. Not. R. Astron. Soc., 357, 401-419 (2005) An XMM-Newton view of M101 - II. Global X-ray source properties.

J/MNRAS/357/471/ 3hed.dat 3heh.dat 4hed.dat 4heh.dat (686Kb) Engel E.A., Doss N., Harris G.J., Tennyson J. Mon. Not. R. Astron. Soc., 357, 471-477 (2005) Calculated spectra for HeH^{+} and its effect on the opacity of cool metal-poor stars.

J/MNRAS/357/497/ refs.dat table1.dat table2.dat table5.dat (50Kb) Bilir S., Karatas Y., Demircan O., Eker Z.

Mon. Not. R. Astron. Soc., 357, 497-517 (2005) Kinematics of W Ursae Majoris type binaries and evidence of the two types of formation.

J/MNRAS/357/645/ table2.dat (43Kb) Almoznino E., Brosch N., Shara M., Zurek D. Mon. Not. R. Astron. Soc., 357, 645-655 (2005) The deepest Hubble Space Telescope far-ultraviolet observations in the Large Magellanic Cloud.

J/MNRAS/357/732/ table1.dat (47Kb) Larios G.R., Phillips J.P. Mon. Not. R. Astron. Soc., 357, 732-752 (2005) An analysis of 2MASS near-infrared photometry for galactic planetary nebulae.

J/MNRAS/357/819/ table3.dat (25Kb) Sabatini S., Davies J., Van Driel W., Baes M., Roberts S., Smith R., Linder S., O’neil K. Mon. Not. R. Astron. Soc., 357, 819-833 (2005) The dwarf low surface brightness galaxy population of the Virgo Cluster - II. Colours and HI line observations.

J/MNRAS/357/1231/ table4.dat table5.dat table6.dat (28Kb) Brand K., Rawlings S., Hill G.J., Tufts J.R. Mon. Not. R. Astron. Soc., 357, 1231-1254 (2005) The three-dimensional clustering of radio galaxies in the Texas-Oxford NVSS structure survey.

J/MNRAS/357/1267/ table1.dat (24Kb) Cirasuolo M., Magliocchetti M., Celotti A. Mon. Not. R. Astron. Soc., 357, 1267-1280 (2005) Faint radio-loud quasars: clues to their evolution.

J/MNRAS/357/1337/ notes.dat table1.dat table2.dat (31Kb) Moss C., Whittle M. Mon. Not. R. Astron. Soc., 357, 1337-1347 (2005) An Hαsurvey of cluster galaxies - V. Cluster-field comparison for early-type galaxies.

J/MNRAS/357/1357/ table2.dat (8Kb) Nakata F., Kodama T., Shimasaku K., Doi M., Furusawa H., Hamabe M., Kimura M., Komiyama Y., Miyazaki S., Okamura S., Ouchi M., Sekiguchi M., Ueda Y., Yagi M., Yasuda N. Mon. Not. R. Astron. Soc., 357, 1357-1362 (2005) Discovery of a large-scale clumpy structure around the Lynx supercluster at z 1.27.

J/MNRAS/358/13/  table1.dat  (14Kb)

Jeffries R.D., Oliveira J.M. Mon. Not. R. Astron. Soc., 358, 13-29 (2005) The Lithium depletion boundary in NGC 2547 as a test of pre-main-sequence evolutionary models.

J/MNRAS/358/30/  table1.dat  (26Kb)

Eyer L., Blake C. Mon. Not. R. Astron. Soc., 358, 30-38 (2005) Automated classification of variable stars for All-Sky Automated Survey 1-2 data.

J/MNRAS/358/66/  table3.dat  (39Kb)

Fabregat J., Capilla G. Mon. Not. R. Astron. Soc., 358, 66-70 (2005) CCD uvbyβphotometry of the young open cluster NGC 663.

J/MNRAS/358/149/ tablea1.dat tablea2.dat (12Kb) Pope A., Borys C., Scott D., Conselice C., Dickinson M., Mobasher B. Mon. Not. R. Astron. Soc., 358, 149-167 (2005) The Hubble Deep Field North SCUBA Super-map. III. Optical and near-infrared properties of submillimetre galaxies.

J/MNRAS/358/233/ tableb1.dat tableb2.dat (11Kb) Fritz A., Ziegler B.L., Bower R.G., Smail I., Davies R.L. Mon. Not. R. Astron. Soc., 358, 233-255 (2005) On the evolutionary status of early-type galaxies in clusters at z 0.2. I. The Fundamental Plane.

J/MNRAS/358/333/ tables1.dat tables2.dat (675Kb) Gonzalez-Solares E.A., Perez-Fournon I., Rowan-Robinson M., Oliver S., Vaccari M., Lari C., Irwin M., Mcmahon R.G., Hodgkin S., Ciliegi P., Serjeant S., Willott C.J. Mon. Not. R. Astron. Soc., 358, 333-340 (2005) The European Large Area ISO Survey: optical identifications of 15-μm and 1.4-GHz sources in N1 and N2.

J/MNRAS/358/341/ ic348.dat table1.dat (30Kb)

Littlefair S.P., Naylor T., Burningham B., Jeffries R.D. Mon. Not. R. Astron. Soc., 358, 341-352 (2005) Do accretion discs regulate the rotation of young stars?

J/MNRAS/358/397/ 15umfa.dat (301Kb) Vaccari M., Lari C., Angeretti L., Fadda D., Gruppioni C., Pozzi F., Prouton O., Aussel H., Babbedge T., Ciliegi P., Franceschini A., Gonzalez-Solares E., Franca F.L., Oliver S., Perez-Fournon I., Rowan-Robinson M., Serjeant S., Vaisanen P. Mon. Not. R. Astron. Soc., 358, 397-418 (2005) Final analysis of ELAIS 15-μm observations: method, reduction and catalogue.

J/MNRAS/358/521/ table1.dat table2.dat table3.dat table4.dat table5.dat (167.1 Mb) Molla M., Diaz A.I. Mon. Not. R. Astron. Soc., 358, 521-543 (2005) A grid of chemical evolution models as a tool to interpret spiral and irregular galaxies data.

J/MNRAS/358/563/ table1.dat (14Kb) Megier A., Krelowski J., Weselak T. Mon. Not. R. Astron. Soc., 358, 563-571 (2005) Correlation patterns between 11 diffuse interstellar bands and ultraviolet extinction.

J/MNRAS/358/795/ tab3-10.dat tab11-13.dat (29Kb) Street R.A., Horne K., Lister T.A., Penny A., Tsapras Y., Quirrenbach A., Safizadeh N., Cooke J., Mitchell D., Cameron A.C. Mon. Not. R. Astron. Soc., 358, 795-812 (2005) Variable stars in the field of open cluster NGC 6819 - II.

J/MNRAS/358/1116/ table1.dat (22Kb) La Barbera F., Covone G., Busarello G., Capaccioli M., Haines C.P., Mercurio A., Merluzzi P. Mon. Not. R. Astron. Soc., 358, 1116-1132 (2005) New insights into the structure of early-type galaxies: the photometric plane at z 0.3.

J/MNRAS/358/1159/ notes.dat tablea1.dat tablea2.dat (78Kb) Muxlow T.W.B., Richards A.M.S., Garrington S.T., Wilkinson P.N., Anderson B., Richards E.A., Axon D.J., Fomalont E.B., Kellermann K.I., Partridge R.B., Windhorst R.A. Mon. Not. R. Astron. Soc., 358, 1159-1194 (2005) High-resolution studies of radio sources in the Hubble Deep and Flanking Fields.

J/MNRAS/358/1215/ table3.dat table4.dat table5.dat table6.dat table7.dat table8.dat table9.dat table10.dat table11.dat table12.dat table13.dat (7.2 Mb) Piatti A.E., Sarajedini A., Geisler D., Seguel J., Clark D. Mon. Not. R. Astron. Soc., 358, 1215-1230 (2005) Tracing the formation history of intermediate-age star clusters in the Small Magellanic Cloud.

J/MNRAS/359/2/  tablea1.dat  (22Kb)

Varricatt W.P., Davis C.J., Adamson A.J. Mon. Not. R. Astron. Soc., 359, 2-15 (2005) Multi-epoch infrared photometric study of the star-forming region G173.58+2.45.

J/MNRAS/359/27 See II/268/

J/MNRAS/359/57 See II/268/

J/MNRAS/359/865/ table3.dat table4.dat table5.dat (55Kb) Yushchenko A., Gopka V., Kim C., Musaev F., Kang Y.W., Kovtyukh V., Soubiran C. Mon. Not. R. Astron. Soc., 359, 865-873 (2005) The chemical composition of δScuti.

J/MNRAS/359/930/ table1.dat (53Kb) Bongiovanni A., Bruzual G., Magris G., Gallego J., Garcia-Dabo C.E., Coppi P., Sabbey C. Mon. Not. R. Astron. Soc., 359, 930-940 (2005) The CIDA-UCM-Yale shallow survey for emission-line galaxies.

J/MNRAS/359/1498/ table1.dat table2.dat table3.dat table4.dat table5.dat (35Kb) Ellingsen S.P. Mon. Not. R. Astron. Soc., 359, 1498-1516 (2005) The relationship between class I and class II methanol masers.

J/MNRAS/359/1524/ table2.dat table3.dat table5.dat (14Kb) Lorimer D.R., Xilouris K.M., Fruchter A.S., Stairs I.H., Camilo F., Vazquez A.M., Eder J.A., McLaughlin M.A., Roberts M.S.E., Hessels J.W.T., ransom S.M. Mon. Not. R. Astron. Soc., 359, 1524-1530 (2005)

Discovery of 10 pulsars in an Arecibo drift-scan survey.

J/MNRAS/360/81 See VII/240/

J/MNRAS/360/185/ catalog.dat (212Kb) Bellazzini M., Gennari N., Ferraro F.R. Mon. Not. R. Astron. Soc., 360, 185-193 (2005) The red giant branch tip and bump of the Leo II dwarf spheroidal galaxy.

J/MNRAS/360/340/ tablea1.dat (9Kb) Cleary K.A., Taylor A.C., Waldram E., Battye R.A., Dickinson C., Davies R.D., Davis R.J., Genova-Santos R., Grainge K., Jones M.E., Kneissl R., Pooley G.G., Rebolo R., Rubino-Martin J.A., Saunders R.D.E., Scott P.F., Slosar A., Titterington D., Watson R.A. Mon. Not. R. Astron. Soc., 360, 340-353 (2005) Source subtraction for the extended Very Small Array and 33-GHz source count estimates.

J/MNRAS/360/587/ table1.dat table2.dat (21Kb) Balogh M.L., Miller C., Nichol R., Zabludoff A., Goto T. Mon. Not. R. Astron. Soc., 360, 587-609 (2005) Near-infrared imaging of 222 nearby Hδ-strong galaxies from the Sloan Digital Sky Survey.

J/MNRAS/360/782/ table1.dat table2.dat table3.dat (16Kb) Georgantopoulos I., Georgakakis A., Koulouridis E. Mon. Not. R. Astron. Soc., 360, 782-790 (2005) The XMM-Newton Needles in the Haystack Survey: the local X-ray luminosity function of ‘normal’ galaxies.

J/MNRAS/360/974/ refs.dat table1.dat table2.dat table3.dat (46Kb) Hobbs G., Lorimer D.R., Lyne A.G., Kramer M. Mon. Not. R. Astron. Soc., 360, 974-992 (2005) A statistical study of 233 pulsar proper motions.

J/MNRAS/360/1305/ table1.dat table2.dat table3.dat table4.dat table5.dat table6.dat (36Kb) Roy S., Rao A.P., Subrahmanyan R. Mon. Not. R. Astron. Soc., 360, 1305-1324 (2005) Extragalactic sources towards the central region of the Galaxy.

J/MNRAS/360/1345/ table1.dat table2.dat (70Kb) Karatas Y., Bilir S., Schuster W.J. Mon. Not. R. Astron. Soc., 360, 1345-1354 (2005) Kinematics, ages and metallicities for F- and G-type stars in the solar neighbourhood.

J/MNRAS/361/109/ table3.dat (15Kb) Bamford S.P., Milvang-Jensen B., Aragon-Salamanca A., Simard L. Mon. Not. R. Astron. Soc., 361, 109-127 (2005) The Tully-Fisher relation of distant cluster galaxies.

J/MNRAS/361/211/ table2.dat (33Kb) Lucas P.W., Roche P.F., Tamura M. Mon. Not. R. Astron. Soc., 361, 211-232 (2005) A deep survey of brown dwarfs in Orion with Gemini.

J/MNRAS/361/272/ ngc6304.dat ngc6569.dat ngc6637.dat ngc6638.dat table13.dat (2.7 Mb) Valenti E., Origlia L., Ferraro, F.R. Mon. Not. R. Astron. Soc. 361, 272 (2005) Near-Infrared photometry of four metal-rich Bulge globular clusters: NGC 6304, NGC 6569, NGC 6637, NGC 6638.

J/MNRAS/361/525/ table2.dat table3.dat table4.dat (15Kb) Doherty M., Bunker A.J., Ellis R.S., McCarthy P.J. Mon. Not. R. Astron. Soc., 361, 525-549 (2005) The Las Campanas Infrared Survey - V. Keck spectroscopy of a large sample of extremely red objects.

J/MNRAS/361/1055/ table1.dat (155Kb) Sabogal B.E., Mennickent R.E., Pietrzynski G., Gieren W. Mon. Not. R. Astron. Soc., 361, 1055-1062 (2005) Be star candidates in the Large Magellanic Cloud: the catalogue and comparison with the Small Magellanic Cloud sample.

J/MNRAS/361/1063/ refs.dat table2.dat (37Kb) Perez-Montero E., Diaz A.I. Mon. Not. R. Astron. Soc., 361, 1063-1076 (2005) A comparative analysis of empirical calibrators for nebular metallicity.

J/MNRAS/361/1091/ sp/* table1.dat table2.dat (6.3 Mb) Naylor T., Allan A., Long K.S. Mon. Not. R. Astron. Soc., 361, 1091-1101 (2005) The masses, radii and luminosities of the components of U Geminorum.

J/MNRAS/361/1102/ table1.dat table2.dat tablea.dat tableb.dat (24Kb) Sim S.A., Jordan C. Mon. Not. R. Astron. Soc., 361, 1102-1120 (2005) Modelling the chromosphere and transition region of ϵEri (K2 V).

J/MNRAS/362/9/  table2.dat  (230Kb)

Best P.N., Kauffmann G., Heckman T.M., Ivezic Z. Mon. Not. R. Astron. Soc., 362, 9-24 (2005) A sample of radio-loud active galactic nuclei in the Sloan Digital Sky Survey.

J/MNRAS/362/117/ table1.dat (9Kb) Kane S.R., Lister T.A., Cameron A.C., Horne K., James D., Pollacco D.L., Street R.A., Tsapras Y. Mon. Not. R. Astron. Soc., 362, 117-126 (2005) Results from the Wide Angle Search for Planets Prototype (WASP0) - II. Stellar variability in the Pegasus field.

J/MNRAS/362/301/ table2.dat (20Kb) Garcia-Rojas J., Esteban C., Peimbert A., Peimbert M., Rodriguez M., Ruiz M.T. Mon. Not. R. Astron. Soc., 362, 301-312 (2005) Deep echelle spectrophotometry of S311, a Galactic HII region located outside the solar circle.

J/MNRAS/362/1006/ tables.dat (17Kb) Todd I., Pollacco D., Skillen I., Bramich D.M., Bell S., Augusteijn T. Mon. Not. R. Astron. Soc., 362, 1006-1014 (2005) A survey of eclipsing binary stars in the eastern spiral arm of M31.

J/MNRAS/362/1025/ (2Kb) Paunzen E., Pintado O.I., Maitzen H.M., Claret A. Mon. Not. R. Astron. Soc., 362, 1025-1030 (2005) On the incidence of chemically peculiar stars in the Large Magellanic Cloud.

J/MNRAS/362/1065/ table1.dat table2.dat table3.dat (17Kb) Wang Q.D., Whitaker K.E., Williams R. Mon. Not. R. Astron. Soc., 362, 1065-1077 (2005) An XMM-Newton and Chandra study of the starburst galaxy IC 10.

J/MNRAS/362/1143/ (3Kb) Mouhcine M., Lewis I., Jones B., Lamareille F., Maddox S.J., Contini T. Mon. Not. R. Astron. Soc., 362, 1143-1156 (2005) The [OII]λ3727/Hαratio of emission-line galaxies in the 2dF Galaxy Redshift Survey.

J/MNRAS/362/1167/ table1.dat tablea.dat (47Kb) Petterson O.K.L., Cottrell P.L., Albrow M.D., Fokin A. Mon. Not. R. Astron. Soc., 362, 1167-1182 (2005) A spectroscopic study of bright southern Cepheids - a high-resolution view of Cepheid atmospheres.

J/MNRAS/362/1189/ refs.dat table1.dat table2.dat table3.dat (33Kb) Zou W.Z., Hobbs G., Wang N., Manchester R.N., Wu X.J., Wang H.X. Mon. Not. R. Astron. Soc., 362, 1189-1198 (2005) Timing measurements and proper motions of 74 pulsars using the Nanshan radio telescope.

J/MNRAS/362/1371/ table5.dat table9.dat (53Kb) Loaring N.S., Dwelly T., Page M.J., Mason K., Mchardy I., Gunn K., Moss D., Seymour N., Newsam A.M., Takata T., Sekguchi K., Sasseen T., Cordova F. Mon. Not. R. Astron. Soc., 362, 1371-1395 (2005) XMM-Newton 13^{H} deep field - I. X-ray sources.

J/MNRAS/363/405/ table5.dat (51Kb) Hill T., Burton M.G., Minier V., Thompson M.A., Walsh A.J., Hunt-Cunningham M., Garay G. Mon. Not. R. Astron. Soc., 363, 405-451 (2005) Millimetre continuum observations of southern massive star formation regions. I. SIMBA observations of cold cores.

J/MNRAS/363/529/ bjohn.dat bstro.dat minima.dat rv.dat ujohn.dat ustro.dat vjohn.dat vstro.dat ystro.dat (182Kb) Southworth J., Smalley B., Maxted P.F.L., Claret A., Etzel P.B.

Mon. Not. R. Astron. Soc., 363, 529-542 (2005) Absolute dimensions of detached eclipsing binaries - I. The metallic-lined system WW Aurigae.

J/MNRAS/363/734/ table1.dat table2.dat tableb.dat (30Kb) Clementini G., Ripepi V., Bragaglia A., Fiorenzano A.F.M., Held E.V., Gratton R.G. Mon. Not. R. Astron. Soc., 363, 734-748 (2005) The metal abundance distribution of the oldest stellar component in the Sculptor dwarf spheroidal galaxy.

J/MNRAS/363/769/ tablea1.dat tablea2.dat tablea3.dat tablea4.dat tablea5.dat tablea6.dat tableb1.dat tableb2.dat tableb3.dat tableb4.dat tableb5.dat tableb6.dat tableb7.dat (50Kb) Samurovic S., Danziger I.J. Mon. Not. R. Astron. Soc., 363, 769-800 (2005) Dark matter in early-type galaxies: dynamical modelling of IC 1459, IC 3370, NGC 3379 and NGC 4105.

J/MNRAS/363/1019/ tablea1.dat tablea2.dat tableb1.dat tableb2.dat tableb3.dat tableb4.dat tablec1.dat (73Kb) Grant N.I., Kuipers J.A., Phillipps S. Mon. Not. R. Astron. Soc., 363, 1019-1030 (2005) Nucleated dwarf elliptical galaxies in the Virgo cluster.

J/MNRAS/363/1111/ gp.dat irasgap.dat lmc.dat sfr.dat smc.dat (309Kb) Clarke A.J., Oudmaijer R.D., Lumsden S.L. Mon. Not. R. Astron. Soc., 363, 1111-1124 (2005) Tycho-2 stars with infrared excess in the MSX Point Source Catalogue.

J/MNRAS/364/117/ table2.dat (180Kb) Matsunaga N., Fukushi H., Nakada Y. Mon. Not. R. Astron. Soc., 364, 117-125 (2005) Mira variables in the Galactic bulge with OGLE-II data.

J/MNRAS/364/146/ tablea1.dat (10Kb) Grutzbauch R., Annibali F., Bressan A., Focardi P., Kelm B., Rampazzo R., Zeilinger W.W. Mon. Not. R. Astron. Soc., 364, 146-162 (2005) Optical properties of the NGC 5328 group of galaxies.

J/MNRAS/364/179/ al1.dat ber80.dat czern37.dat king25.dat lynga9.dat ngc5764.dat vdbh150.dat vdbh261.dat (1.1 Mb) Carraro G., Janes K.A., Eastman J.D. Mon. Not. R. Astron. Soc., 364, 179-186 (2005) Photometry of neglected open clusters in the first and fourth Galactic quadrants.

J/MNRAS/364/272/ table1.dat table2.dat table3.dat table4.dat (45Kb) Ford A., Jeffries R.D., Smalley B. Mon. Not. R. Astron. Soc., 364, 272-282 (2005) Elemental abundances in the Blanco 1 open cluster.

J/MNRAS/365/110/ h10cz29.dat ngc2425.dat (584Kb) Pietrukowicz P., Kaluzny J., Krzeminski W. Mon. Not. R. Astron. Soc. 365, 110 (2006) CCD photometry of distant open clusters NGC 2425, Haffner 10 and Czernik 29.

J/MNRAS/365/401/ table3.dat table4.dat (19Kb) Reid W.A., Parker Q.A. Mon. Not. R. Astron. Soc., 365, 401-413 (2006) A new population of planetary nebulae discovered in the Large Magellanic Cloud - I. Preliminary sample.

J/MNRAS/365/439/ table2.dat (14Kb) Christopher N.M., Smail I. Mon. Not. R. Astron. Soc., 365, 439-446 (2006) A catalogue of potential adaptive optics survey fields from the UKIRT archive.

J/MNRAS/365/447/ table1.dat table2.dat (18Kb) Casewell S.L., Jameson R.F., Dobbie P.D. Mon. Not. R. Astron. Soc., 365, 447-453 (2006) New stellar members of the Coma Berenices open star cluster.

J/MNRAS/365/454/ refs.dat table1.dat (52Kb) Hoyos C., Diaz A.I. Mon. Not. R. Astron. Soc., 365, 454-468 (2006) The impact of the visibility of the [OIII]λ4363 line on the general properties of HII galaxies in the Local Universe.

J/MNRAS/365/1357/ table1.dat table1.ps (62Kb) Criscienzo M.D., Caputo F., Marconi M., Musella I. Mon. Not. R. Astron. Soc., 365, 1357-1366 (2006) RR Lyrae-based calibration of the Globular Cluster Luminosity Function.

J/MNRAS/365/1392/ table2.dat (21Kb) Tanaka M., Kodama T., Arimoto N., Tanaka I. Mon. Not. R. Astron. Soc., 365, 1392-1404 (2006) Spectroscopically confirmed large-scale structures associated to a z=0.83 cluster.

J/MNRAS/366/144/ table3.dat (14Kb) Nakamura O., Aragon-Salamanca A., Milvang-Jensen B., Arimoto N., Ikuta C., Bamford S.P. Mon. Not. R. Astron. Soc., 366, 144-162 (2006) The Tully-Fisher relation of intermediate redshift field and cluster galaxies from Subaru spectroscopy.

J/MNRAS/366/171/ table1.dat (15Kb) Georgakakis A.E., Georgantopoulos I., Akylas A. Mon. Not. R. Astron. Soc., 366, 171-181 (2006) Red active galactic nuclei in XMM-Newton/Sloan Digital Sky Survey fields.

J/MNRAS/366/267/ table1.dat table2.dat (61Kb) Fairall A.P., Woudt P.A. Mon. Not. R. Astron. Soc., 366, 267-273 (2006) A sample of galaxies near the South Celestial Pole.

J/MNRAS/366/480/ table1.dat (12Kb) Gu Q., Melnick J., Fernandes R.C., Kunth D., Terlevich E., Terlevich R. Mon. Not. R. Astron. Soc., 366, 480-490 (2006) Emission-line properties of Seyfert 2 nuclei.

J/MNRAS/366/645/ larcs3.dat (484Kb) Pimbblet K.A., Smail I., Edge A.C., O’hely E., Couch W.J., Zabludoff A.I. Mon. Not. R. Astron. Soc., 366, 645-666 (2006) The Las Campanas/Anglo-Australian Telescope Rich Cluster Survey - III. Spectroscopic studies of X-ray bright galaxy clusters at z 0.1.

J/MNRAS/366/739/ table4.dat (65Kb) Arias J.I., Barba R.H., Maiz-Apellaniz J., Morrell N.I., Rubio M. Mon. Not. R. Astron. Soc., 366, 739-757 (2006) The infrared Hourglass cluster in M8.

J/MNRAS/366/1096/ tablea1.dat tablea2.dat (22Kb) Busfield A.L., Purcell C.R., Hoare M.G., Lumsden S.L., Moore T.J.T., Oudmaijer R.D. Mon. Not. R. Astron. Soc., 366, 1096-1117 (2006) Resolving the kinematic distance ambiguity of southern massive young stellar object candidates.

J/MNRAS/366/1230/ tablea1.dat tablea2.dat tablea3.dat tablea4.dat (86Kb) Forbes D.A., Sanchez-Blazquez P., Phan A.T.T., Brodie J.P., Strader J., Spitler L. Mon. Not. R. Astron. Soc., 366, 1230-1242 (2006) An imaging study of the globular cluster systems of NGC 1407 and 1400.

J/MNRAS/366/1265/ notes.dat table1.dat table2.dat table4.dat (37Kb) Brookes M.H., Best P.N., Rengelink R., Roettgering H.J.A. Mon. Not. R. Astron. Soc., 366, 1265-1288 (2006) CENSORS: A Combined EIS-NVSS Survey of Radio Sources - II. Infrared imaging and the K-z relation.

J/MNRAS/367/156/ table2.dat (136Kb) Bassino L.P., Richtler T., Dirsch B. Mon. Not. R. Astron. Soc., 367, 156-168 (2006) Globular cluster systems in low-luminosity early-type galaxies near the Fornax cluster centre.

J/MNRAS/367/400 See VI/121/

J/MNRAS/367/513/ table2.dat (13Kb) Tapia M., Persi P., Bohigas J., Roth M., Gomez M. Mon. Not. R. Astron. Soc., 367, 513-526 (2006) Imaging study of NGC 3372, the Carina nebula - II. Evidence of activity in the complex Trumpler 14/Car I photodissociation region.

J/MNRAS/367/599/ table1.dat (1.2 Mb) Piatti A.E., Claria J.J., Ahumada A.V. Mon. Not. R. Astron. Soc., 367, 599-610 (2006)

First estimates of the fundamental parameters of the relatively bright Galactic open cluster NGC 5288.

J/MNRAS/367/627/ tablea1.dat (37Kb) Ellis S.C., O’sullivan E. Mon. Not. R. Astron. Soc., 367, 627-645 (2006) Correlations of near-infrared, optical and X-ray luminosity for early-type galaxies.

J/MNRAS/367/781/ table1.dat table2.dat table3.dat table7.dat (79Kb) Jeffries R.D., Evans P.A., Pye J.P., Briggs K.R. Mon. Not. R. Astron. Soc. 367, 781 (2006) An XMM-Newton observation of the young open cluster NGC 2547: coronal activity at 30 Myr.

J/MNRAS/367/1170/ table1.dat (21Kb) Scardia M., Prieur J.-L., Pansecchi L., Argyle R.W., Sala M., Ghigo M., Koechlin L., Aristidi E. Mon. Not. R. Astron. Soc., 367, 1170-1180 (2006) Speckle observations with PISCO in Merate - II. Astrometric measurements of visual binaries in 2004.

J/MNRAS/367/1261/ tab3-18.dat table1.dat (57Kb) Morris S.L., Jannuzi B.T. Mon. Not. R. Astron. Soc., 367, 1261-1281 (2006) The association between gas and galaxies - I. CFHT spectroscopy and pair analysis.

J/MNRAS/367/1329/ table2.dat table45.dat (54Kb) Reddy B.E., Lambert D.L., Prieto C.A. Mon. Not. R. Astron. Soc., 367, 1329-1366 (2006) Elemental abundance survey of the Galactic thick disc.

J/MNRAS/367/1441/ table1.dat (491Kb) Baume G., Moitinho A., Vazquez R.A., Solivella G., Carraro G., Villanova S. Mon. Not. R. Astron. Soc., 367, 1441-1449 (2006) NGC 2401: a template of the young population of the Norma-Cygnus arm in the Third Galactic Quadrant.

J/MNRAS/367/1478/ table2.dat table3.dat table4.dat (134Kb) Hunter I., Smoker J.V., Keenan F.P., Ledoux C., Jehin E., Cabanac R., Melo C., Bagnulo S. Mon. Not. R. Astron. Soc., 367, 1478-1514 (2006) Early-type stars observed in the ESO UVES Paranal Observatory Project - I. Interstellar NaI UV, TiII and CaII K observations.

J/MNRAS/367/1609/ table1.dat table2.dat table3.dat (21Kb) Bains I., Wong T., Cunningham M., Sparks P., Brisbin D., Calisse P., Dempsey J.T., Deragopian G., Ellingsen S., Fulton B., Herpin F., Jones P., Kouba Y., Kramer C., Ladd E.F., Longmore S.N., McEvoy J., Maller M., Minier V., Mookerjea B., Phillips C., Purcell C.R., Walsh A., Voronkov M.A., Burton M.G. Mon. Not. R. Astron. Soc., 367, 1609-1628 (2006) Molecular line mapping of the giant molecular cloud associated with RCW 106. I. ^{13}CO.

J/MNRAS/368/247/ refs.dat table1.dat table2.dat tablea1.dat tablea2.dat tablea3.dat tablea4.dat tablea5.dat tablea6.dat (28Kb) Catanzaro G. Mon. Not. R. Astron. Soc., 368, 247-252 (2006) Spectroscopic analyses of six suspected chemically peculiar stars.

J/MNRAS/368/253/ table2.dat (60Kb) Garcia-Rojas J., Esteban C., Peimbert M., Costado M.T., Rodriguez M., Peimbert A., Ruiz M.T. Mon. Not. R. Astron. Soc., 368, 253-279 (2006) Faint emission lines in the Galactic HII regions M16, M20 and NGC 3603.

J/MNRAS/368/534/ table1.dat (12Kb) Nagayama T., Woudt P.A., Wakamatsu K., Nishiyama S., Nagashima C., Kato D., Nagata T., Nakaya H., Sugitani K., Tamura M., Sato S. Mon. Not. R. Astron. Soc., 368, 534-543 (2006) Near-infrared study of CIZA J1324.7-5736, the second richest cluster of galaxies in the Great Attractor.

J/MNRAS/368/651/ catalog.dat (26Kb) Pimbblet K.A., Roseboom I.G., Doyle M.T. Mon. Not. R. Astron. Soc., 368, 651-658 (2006) The large peculiar velocity of the cD galaxy in Abell 3653.

J/MNRAS/368/1087 See VI/119/

J/MNRAS/368/1223/ table3.dat table4.dat (45Kb) Hill T., Thompson M.A., Burton M.G., Walsh A.J., Minier V., Cunningham M.R., Pierce-Price D. Mon. Not. R. Astron. Soc., 368, 1223-1268 (2006)

Millimetre continuum observations of southern massive star formation regions - II. SCUBA observations of cold cores and the dust grain emissivity index (β).

J/MNRAS/368/1392/ table2.dat (10Kb) Pettersen B.R. Mon. Not. R. Astron. Soc., 368, 1392-1394 (2006) Flare variability in the close binary FL Vir.

J/MNRAS/368/1411/ refs.dat table4.dat table5.dat (153Kb) Augusto P., Gonzalez-Serrano J.I., Perez-Fournon I., Wilkinson P.N. Mon. Not. R. Astron. Soc. 368, 1411 (2006) Flat-spectrum symmetric objects with 1kpc sizes. I. The candidates.

J/MNRAS/368/1757/ table1.dat table2.dat (73Kb) Wils P., Lloyd C., Bernhard K. Mon. Not. R. Astron. Soc. 368, 1757 (2006) A catalogue of RR Lyrae stars from the Northern Sky Variability Survey.

J/MNRAS/368/1959/ table2.dat table4.dat (24Kb) Liu X.-W., Barlow M.J., Zhang Y., Bastin R.J., Storey P.J. Mon. Not. R. Astron. Soc., 368, 1959-1970 (2006) Chemical abundances for Hf 2-2, a planetary nebula with the strongest-known heavy-element recombination lines.

J/MNRAS/369/97/  table1.dat tablea1.dat  (63Kb)

Halliday C., Carter D., Bridges T.J., Jackson Z.C., Wilkinson M.I., Quinn D.P., Evans N.W., Douglas N.G., Merrett H.R., Merrifield M.R., Romanowsky A.J., Kuijken K., Irwin M.J. Mon. Not. R. Astron. Soc., 369, 97-119 (2006) Planetary nebula velocities in the disc and bulge of M31.

J/MNRAS/369/120/ table3.dat (330Kb) Merrett H.R., Merrifield M.R., Douglas N.G., Kuijken K., Romanowsky A.J., Napolitano N.R., Arnaboldi M., Capaccioli M., Freeman K.C., Gerhard O., Coccato L., Carter D., Evans N.W., Wilkinson M.I., Halliday C., Bridges T.J. Mon. Not. R. Astron. Soc., 369, 120-142 (2006) A deep kinematic survey of planetary nebulae in the Andromeda galaxy using the Planetary Nebula Spectrograph.

J/MNRAS/369/156/ table1.dat (11Kb) Page M.J., Loaring N.S., Dwelly T., Mason K.O., McHardy I., Gunn K., Moss D., Sasseen T., Cordova F., Kennea J., Seymour N. Mon. Not. R. Astron. Soc., 369, 156-170 (2006) X-ray spectra of sources in the 13^{H} XMM-Newton/Chandra deep field.

J/MNRAS/369/449/ table2.dat table3.dat (27Kb) Kedziora-Chudczer L. Mon. Not. R. Astron. Soc., 369, 449-464 (2006) Long-term monitoring of the intra-day variable quasar PKS 0405-385.

J/MNRAS/369/697/ refs.dat table1.dat tablea1.dat (21Kb) Goudfrooij P., Gilmore D., Kissler-Patig M., Maraston C. Mon. Not. R. Astron. Soc., 369, 697-704 (2006) Integrated-light VRI imaging photometry of globular clusters in the Magellanic Clouds.

J/MNRAS/369/723/ tablea1.dat tablea2.dat tablea3.dat tablea4.dat tablea5.dat tablea6.dat (302Kb) Ngeow C., Kanbur S.M Mon. Not. R. Astron. Soc. 369, 723 (2006) Period-colour and amplitude-colour relations in classical Cepheid variables. IV: The multi-phase relations

J/MNRAS/369/751/ notes.dat table1.dat table2.dat table3.dat table6.dat (322Kb) Whitelock P.A., Feast M.W., Marang F., Groenewegen M.A.T. Mon. Not. R. Astron. Soc., 369, 751-782 (2006) Near-infrared photometry of carbon stars.

J/MNRAS/369/783/ table2.dat table3.dat table4.dat (26Kb) Menzies J.W., Feast M.W., Whitelock P.A. Mon. Not. R. Astron. Soc., 369, 783-790 (2006) Carbon-rich Mira variables: radial velocities and distances.

J/MNRAS/369/1131/ table2.dat (206Kb) Radburn-Smith D.J., Lucey J.R., Woudt P.A., Kraan-Korteweg R.C., Watson F.G.

Mon. Not. R. Astron. Soc. 369, 1131 (2006) Structures in the Great Attractor Region.

J/MNRAS/369/1334/ catalog.dat cluster/* table4.dat (8.2 Mb) Sochting I.K., Huber M.E., Clowes R.G., Howell S.B. Mon. Not. R. Astron. Soc., 369, 1334-1350 (2006) The FSVS Cluster Catalogue: galaxy clusters and groups in the Faint Sky Variability Survey.

J/MNRAS/369/1351/ tablea1.dat tableb1.dat (56Kb) Brough S., Forbes D.A., Kilborn V.A., Couch W., Colless M. Mon. Not. R. Astron. Soc., 369, 1351-1374 (2006) Eridanus - a supergroup in the local Universe?

J/MNRAS/369/1375/ table2.dat (31Kb) Trentham N., Tully R.B., Mahdavi A. Mon. Not. R. Astron. Soc., 369, 1375-1391 (2006) Dwarf galaxies in the dynamically evolved NGC 1407 Group.

J/MNRAS/369/1677/ tablea1.dat (12Kb) Pandey G., Reddy B.E. Mon. Not. R. Astron. Soc., 369, 1677-1682 (2006) Abundance analysis of the cool extreme helium star LSS 3378.

J/MNRAS/369/1822/ table1.dat table3.dat (28Kb) Maheswar G., Bhatt H.C. Mon. Not. R. Astron. Soc., 369, 1822-1836 (2006) Photometric distances to nine dark globules.

J/MNRAS/369/1880/ table1.dat table2.dat table3.dat table4.dat table5.dat table6.dat table7.dat (31Kb) Elias-Rosa N., Benetti S., Cappellaro E., Turatto M., Mazzali P.A., Patat F., Meikle W.P.S., Stehle M., Pastorello A., Pignata G., Kotak R., Harutyunyan A., Altavilla G., Navasardyan H., Qiu Y., Salvo M., Hillebrandt W. Mon. Not. R. Astron. Soc., 369, 1880-1900 (2006) Anomalous extinction behaviour towards the Type Ia SN 2003cg.

J/MNRAS/370/25/  table2.dat  (25Kb)

Stobbart A.-M., Roberts T.P., Warwick R.S. Mon. Not. R. Astron. Soc., 370, 25-42 (2006) The X-ray properties of the dwarf Magellanic-type galaxy NGC 55.

J/MNRAS/370/163/ table2.dat table3.dat (28Kb) Bond J.C., Tinney C.G., Butler R.P., Jones H.R.A., Marcy G.W., Penny A.J., Carter B.D. Mon. Not. R. Astron. Soc. 370, 163 (2006) The abundance distribution of stars with planets.

J/MNRAS/370/435/ photom.dat table1.dat table2.dat (3.5 Mb) Kozlowski S., Wozniak P.R., Mao S., Smith M.C., Sumi T., Vestrand W.T., Wyrzykowski L. Mon. Not. R. Astron. Soc. 370, 435 (2006) Mapping stellar kinematics across the Galactic bar: HST measurements of proper motions in 35 fields.

J/MNRAS/370/954/ tablea1.dat (17Kb) Irwin J., Aigrain S., Hodgkin S., Irwin M., Bouvier J., Clarke C., Hebb L., Moraux E. Mon. Not. R. Astron. Soc., 370, 954-974 (2006) The Monitor project: rotation of low-mass stars in the open cluster M34.

J/MNRAS/370/1034/ table2.dat (15Kb) Carballo R., Gonzalez-Serrano J.I., Montenegro-Montes F.M., Benn C.R., Mack K.-H., Pedani M., Vigotti M. Mon. Not. R. Astron. Soc., 370, 1034-1045 (2006) A FIRST-APM-SDSS survey for high-redshift radio QSOs.

J/MNRAS/370/1057/ tables.dat (27Kb) Scott S.E., Dunlop J.S., Serjeant S. Mon. Not. R. Astron. Soc., 370, 1057-1105 (2006) A combined re-analysis of existing blank-field SCUBA surveys: comparative 850-μm source lists, combined number counts, and evidence for strong clustering of the bright submillimetre galaxy population on arcminute scales.

J/MNRAS/370/1181/ tablea1.dat tablea2.dat tablea3.dat tableb1.dat tableb2.dat tableb3.dat (90Kb) Zhang F., Li L. Mon. Not. R. Astron. Soc., 370, 1181-1184 (2006) A library of Lick Observatory Image Dissector Scanner indices for binary stellar populations.

J/MNRAS/370/1223/ table3.dat tablea1.dat (37Kb) Brough S., Forbes D.A., Kilborn V.A., Couch W. Mon. Not. R. Astron. Soc. 370, 1223 (2006) Southern GEMS groups. I. Dynamical properties.

J/MNRAS/370/1429/ table11.dat (29Kb) Shporer A., Mazeh T. Mon. Not. R. Astron. Soc. 370, 1429 (2006) Long-term V-band monitoring of the bright stars of M33 at the Wise Observatory.

J/MNRAS/370/1979/ table2.dat table3.dat (65Kb) Matsunaga N., Fukushi H., Nakada Y., Tanabe T., Feast M.W., Menzies J.W., Ita Y., Nishiyama S., Baba D., Naoi T., Nakaya H., Kawadu T., Ishihara A., Kato D. Mon. Not. R. Astron. Soc., 370, 1979-1990 (2006) The period-luminosity relation for type II Cepheids in globular clusters.

J/MNRAS/370/2013/ refs.dat table1.dat table2.dat table3.dat table4.dat table5.dat (38Kb) Soydugan E., Soydugan F., Demircan O., Ibanoglu C. Mon. Not. R. Astron. Soc., 370, 2013-2024 (2006) A catalogue of close binaries located in the δScuti region of the Cepheid instability strip.

J/MNRAS/370/2091/ tablea1.dat (21Kb) Chen Y.Q., Zhao G. Mon. Not. R. Astron. Soc., 370, 2091-2096 (2006) The puzzling abundance pattern of HD 134439 and HD 134440.

J/MNRAS/371/38/  table1.dat  (22Kb)

Wang Q.D., Dong H., Lang C. Mon. Not. R. Astron. Soc., 371, 38-54 (2006) The interplay between star formation and the nuclear environment of our Galaxy: deep X-ray observations of the Galactic centre Arches and Quintuplet clusters.

J/MNRAS/371/147/ table2.dat (10Kb) Li Z., Wang Q.D., Irwin J.A., Chaves T. Mon. Not. R. Astron. Soc., 371, 147-156 (2006) An XMM-Newton observation of the massive edge-on Sb galaxy NGC 2613.

J/MNRAS/371/221/ table2.dat (25Kb) Georgakakis A., Nandra K., Laird E.S., Gwyn S., Steidel C.C., Sarajedini V.L., Barmby P., Faber S.M., Coil A.L., Cooper M.C., Davis M., Newman J.A. Mon. Not. R. Astron. Soc., 371, 221-234 (2006) A deep Chandra survey of the Groth Strip. II. Optical identification of the X-ray sources.

J/MNRAS/371/252/ table1.dat table2.dat (32Kb) Levenhagen R.S., Leister N.V. Mon. Not. R. Astron. Soc., 371, 252-262 (2006) Spectroscopic analysis of southern B and Be stars.

J/MNRAS/371/495/ catalog.dat (1.5 Mb) Bouche N., Murphy M.T., Peroux C., Csabai I., Wild V. Mon. Not. R. Astron. Soc., 371, 495-512 (2006) New perspectives on strong z= 0.5 MgII absorbers: are halo mass and equivalent width anticorrelated ?

J/MNRAS/371/703/ catalog.dat sp/* (15.8 Mb) Sanchez-Blazquez P., Peletier R.F., Jimenez-Vicente J., Cardiel N., Cenarro A.J., Falcon-Barroso J., Gorgas J., Selam S., Vazdekis A. Mon. Not. R. Astron. Soc., 371, 703-718 (2006) Medium-resolution Isaac Newton Telescope library of empirical spectra.

J/MNRAS/371/879/ table1.dat (10Kb) Kovtyukh V.V., Soubiran C., Bienayme O., Mishenina T.V., Belik S.I. Mon. Not. R. Astron. Soc., 371, 879-884 (2006) High-precision effective temperatures of 215 FGK giants from line-depth ratios.

J/MNRAS/371/898/ table3.dat (31Kb) Sadler E.M., Ricci R., Ekers R.D., Ekers J.A., Hancock P.J., Jackson C.A., Kesteven M.J., Murphy T., Phillips C., Reinfrank R.F., Staveley-Smith L., Subrahmanyan R., Walker M.A., Wilson W.E., De Zotti G. Mon. Not. R. Astron. Soc., 371, 898-914 (2006) The properties of extragalactic radio sources selected at 20GHz.

J/MNRAS/371/963/ elais.dat hdfn.dat lockman.dat (109Kb)

Biggs A.D., Ivison R.J. Mon. Not. R. Astron. Soc., 371, 963-971 (2006) A catalogue of μJy radio sources in northern legacy fields.

J/MNRAS/371/1140/ table10.dat table11.dat (50Kb) Griffin R.F., Cornell A.P. Mon. Not. R. Astron. Soc., 371, 1140-1158 (2006) Photoelectric radial velocities. Paper XVI: 625 ninth-magnitude K0 stars in the six southern Clube Selected Areas.

J/MNRAS/371/1159/ table1.dat table31.dat (14Kb) Griffin R.F. Mon. Not. R. Astron. Soc., 371, 1159-1172 (2006) Photoelectric radial velocities. Paper XVII: The orbits of 30 spectroscopic binaries in the southern Clube Selected Areas.

J/MNRAS/371/1301/ ber5.dat ber44.dat ber52.dat ber56.dat ngc6827.dat skiff1.dat (3.6 Mb) Carraro G., Subramaniam A., Janes K.A. Mon. Not. R. Astron. Soc., 371, 1301-1310 (2006) Fundamental parameters of six neglected old open clusters.

J/MNRAS/371/1405/ table24.dat table56.dat (18.6 Mb) Hartman J.D., Bersier D., Stanek K.Z., Beaulieu J.-P., Kaluzny J., Marquette J.-B., Stetson P.B., Schwarzenberg-Czerny A. Mon. Not. R. Astron. Soc., 371, 1405-1417 (2006) Deep Canada-France-Hawaii Telescope photometric survey of the entire M33 galaxy - I. Catalogue of 36000 variable point sources.

J/MNRAS/371/1744/ tables1.dat tables2.dat tables3.dat tables4.dat tables5.dat tables6.dat tables7.dat tables8.dat tables9.dat tables10.dat (669Kb) Pitman K.M., Speck A.K., Hofmeister A.M. Mon. Not. R. Astron. Soc., 371, 1744-1754 (2006) Challenging the identification of nitride dust in extreme carbon star spectra.

J/MNRAS/371/1777/ notes.dat tabb456.dat table12.dat tablea1.dat tablea2.dat tablea3.dat tableb1.dat tableb2.dat tableb3.dat (67Kb) Johnson O., Best P., Zaritsky D., Clowe D., Aragon-Salamanca A., Halliday C., Jablonka P., Milvang-Jensen B., Pello R., Poggianti B.M., Rudnick G., Saglia R., Simard L., White S. Mon. Not. R. Astron. Soc. 371, 1777 (2006) The X-ray properties of optically selected z 0.6 clusters in the European Southern Observatory Distant Cluster Survey.

J/MNRAS/371/1793/ refs.dat table1.dat (63Kb) Karatas Y., Schuster W.J. Mon. Not. R. Astron. Soc., 371, 1793-1812 (2006) Metallicity and absolute magnitude calibrations for UBV photometry.

J/MNRAS/371/1891/ table1.dat (65Kb) Rodighiero G., Lari C., Pozzi F., Gruppioni C., Fadda D., Franceschini A., Lonsdale C., Surace J., Shupe D., Fang F. Mon. Not. R. Astron. Soc., 371, 1891-1897 (2006) Counting individual galaxies from deep 24-μm Spitzer surveys.

J/MNRAS/372/163/ table2.dat (24Kb) Jenkins J.S., Jones H.R.A., Tinney C.G., Butler R.P., McCarthy C., Marcy G.W., Pinfield D.J., Carter B.D., Penny A.J. Mon. Not. R. Astron. Soc., 372, 163-173 (2006) An activity catalogue of southern stars.

J/MNRAS/372/425/ catalog.dat (3.4 Mb) Cannon R., Drinkwater M., Edge A., Eisenstein D., Nichol R., Outram P., Pimbblet K., De Propris R., Roseboom I., Wake D., Allen P., Bland-Hawthorn J., Bridges T., Carson D., Chiu K., Colless M., Couch W., Croom S., Driver S., Fine S., Hewett P., Loveday J., Ross N., Sadler E.M., Shanks T., Sharp R., Smith J.A., Stoughton C., Weilbacher P., Brunner R.J., Meiksin A., Schneider D.P. Mon. Not. R. Astron. Soc., 372, 425-442 (2006) The 2dF-SDSS LRG and QSO (2SLAQ) Luminous Red Galaxy Survey.

J/MNRAS/372/661/ Xatlas/* obcensus.dat refs.dat (549Kb) Sana H., Rauw G., Naze N., Gosset E., Vreux J.-M. Mon. Not. R. Astron. Soc. 372, 661 (2006)

An XMM-Newton view of the young open cluster NGC 6231. II. The OB star population.

J/MNRAS/372/741/ table3.dat (91Kb) Simpson C., Martinez-Sansigre A., Rawlings S., Ivison R., Akiyama M., Sekiguchi K., Takata T., Ueda Y., Watson M. Mon. Not. R. Astron. Soc., 372, 741-757 (2006) Radio imaging of the Subaru/XMM-Newton Deep Field. I. The 100-μJy catalogue, optical identifications, and the nature of the faint radio source population.

J/MNRAS/372/777/ table1.dat table2.dat table3.dat tables1.dat tables2.dat (1.0 Mb) Lorimer D.R., Faulkner A.J., Lyne A.G., Manchester R.N., Kramer M., Mclaughlin M.A., Hobbs G., Possenti A., Stairs I.H., Camilo F., Burgay M., D’Amico N., Corongiu A., Crawford F. Mon. Not. R. Astron. Soc., 372, 777-800 (2006) The Parkes Multibeam Pulsar Survey - VI. Discovery and timing of 142 pulsars and a Galactic population analysis.

J/MNRAS/372/1117/ table3.dat (9Kb) Christian D.J., Pollacco D.L., Skillen I., Street R.A., Keenan F.P., Clarkson W.I., Cameron A.C., Kane S.R., Lister T.A., West R.G., Enoch B., Evans A., Fitzsimmons A., Haswell C.A., Hellier C., Hodgkin S.T., Horne K., Irwin J., Norton A.J., Osborne J., Ryans R., Wheatley P.J., Wilson D.M. Mon. Not. R. Astron. Soc., 372, 1117-1128 (2006) The SuperWASP wide-field exoplanetary transit survey: candidates from fields 23h RA 03h.

J/MNRAS/372/1259/ objects.dat table5.dat table6.dat (17Kb) Sharina M.E., Afanasiev V.L., Puzia T.H. Mon. Not. R. Astron. Soc., 372, 1259-1274 (2006) Ages, metallicities and [α/Fe] ratios of globular clusters in NGC 147, 185 and 205.

J/MNRAS/372/1315/ table2.dat (12Kb) Sahu D.K., Anupama G.C., Srividya S., Muneer S. Mon. Not. R. Astron. Soc., 372, 1315-1324 (2006) Photometric and spectroscopic evolution of the Type IIP supernova SN 2004et.

J/MNRAS/372/1425/ tableb1.dat tablec2.dat (22Kb) Halkola A., Seitz S., Pannella M. Mon. Not. R. Astron. Soc., 372, 1425-1462 (2006) Parametric strong gravitational lensing analysis of Abell 1689.

J/MNRAS/372/1621/ table2.dat (22Kb) Coppin K., Chapin E.L., Mortier A.M.J., Scott S.E., Borys C., Dunlop J.S., Halpern M., Hughes D.H., Pope A., Scott D., Serjeant S., Wagg J., Alexander D.M., Almaini O., Aretxaga I., Babbedge T., Best P.N., Blain A., Chapman S., Clements D.L., Crawford M., Dunne L., Eales S.A., Edge A.C., Farrah D., Gaztanaga E., Gear W.K., Granato G.L., Greve T.R., Fox M., Ivison R.J., Jarvis M.J., Jenness T., Lacey C., Lepage K., Mann R.G., Marsden G., Martinez-Sansigre A., Oliver S., Page M.J., Peacock J.A., Pearson C.P., Percival W.J., Priddey R.S., Rawlings S., Rowan-Robinson M., Savage R.S., Seigar M., Sekiguchi K., Silva L., Simpson C., Smail I., Stevens J.A., Takagi T., Vaccari M., Van Kampen E., Willott C.J. Mon. Not. R. Astron. Soc., 372, 1621-1652 (2006) The SCUBA Half-Degree Extragalactic Survey (SHADES). II. Submillimetre maps, catalogue and number counts.

J/MNRAS/372/1657/ bmag/* imag/* table3.dat vmag/* (4.1 Mb) Benko J.M., Bakos G.A., Nuspl J. Mon. Not. R. Astron. Soc., 372, 1657-1674 (2006) Multicolour CCD photometry of the variable stars in globular cluster M3.

J/MNRAS/373/13/  table1.dat  (19Kb)

Casagrande L., Portinari L., Flynn C. Mon. Not. R. Astron. Soc., 373, 13-44 (2006) Accurate fundamental parameters for lower main-sequence stars.

J/MNRAS/373/79 See V/127/

J/MNRAS/373/95/  table2.dat  (14Kb)

Lodieu N., Hambly N.C., Jameson R.F. Mon. Not. R. Astron. Soc., 373, 95-104 (2006) New members in the Upper Scorpius association from the UKIRT Infrared Deep Sky Survey Early Data Release.

J/MNRAS/373/209/ table2.dat table3.dat table4.dat (55Kb) Hu S.M., Wu J.H., Zhao G., Zhou X. Mon. Not. R. Astron. Soc., 373, 209-216 (2006)

Optical multiband observations of BL Lacertae during the outburst of 2005.

J/MNRAS/373/435/ refs.dat table1.dat table2.dat (27Kb) Ibanoglu C., Soydugan F., Soydugan E., Dervisoglu A. Mon. Not. R. Astron. Soc., 373, 435-448 (2006) Angular momentum evolution of Algol binaries.

J/MNRAS/373/521/ table6.dat table7.dat (67Kb) Reid W.A., Parker Q.A. Mon. Not. R. Astron. Soc., 373, 521-550 (2006) A new population of planetary nebulae discovered in the Large Magellanic Cloud. II. Complete PN catalogue.

J/MNRAS/373/687/ int.dat lt.dat spectra.dat spectra/* (12.7 Mb) Southworth J., Gansicke B.T., Marsh T.R., De Martino D., Hakala P., Littlefair S., Rodriguez-gil P., Szkody P. Mon. Not. R. Astron. Soc., 373, 687-699 (2006) VLT/FORS spectroscopy of faint cataclysmic variables discovered by the Sloan Digital Sky Survey.

J/MNRAS/373/705/ table2.dat table3.dat table5.dat (21Kb) Reyle C., Scholz R.D., Schultheis M., Robin A.C., Irwin M. Mon. Not. R. Astron. Soc. 373, 705 (2006) Optical spectroscopy of high proper motion stars: new M dwarfs within 10pc and the closest pair of subdwarfs.

J/MNRAS/373/781/ refs.dat table1.dat table3.dat (27Kb) Leggett S.K., Currie M.J., Varricatt W.P., Hawarden T.G., Adamson A.J., Buckle J., Carroll T., Davies J.K., Davis C.J., Kerr T.H., Kuhn O.P., Seigar M.S., Wold T. Mon. Not. R. Astron. Soc., 373, 781-792 (2006) JHK observations of faint standard stars in the Mauna Kea Observatories near-infrared photometric system.

J/MNRAS/373/1203/ table1.dat table2.dat table3.dat table4.dat table5.dat (31Kb) Kerton C.R. Mon. Not. R. Astron. Soc. 373, 1203 (2006) A sharper view of the outer Galaxy at 1420 and 408 MHz from the Canadian Galactic Plane Survey. I: Revisiting the KR catalogue and new Gigahertz Peaked Spectrum sources.

J/MNRAS/373/1483/ table1.dat table2.dat (25Kb) Eker Z., Demircan O., Bilir S., Karatas Y. Mon. Not. R. Astron. Soc., 373, 1483-1494 (2006) Dynamical evolution of active detached binaries on the logJo-logM diagram and contact binary formation.

J/MNRAS/373/1531/ tablea1.dat (11Kb) Cruz M.J., Jarvis M.J., Blundell K.M., Rawlings S., Croft S., Klockner H.-R., McLure R.J., Simpson C., Targett T.A., Willott C.J. Mon. Not. R. Astron. Soc., 373, 1531-1562 (2006) The 6C** sample of steep-spectrum radio sources - I. Radio data, near-infrared imaging and optical spectroscopy.

J/MNRAS/374/372/ table5.dat tableb1.dat tablec1.dat (25Kb) Lodieu N., Hambly N.C., Jameson R.F., Hodgkin S.T., Carraro G., Kendall T.R. Mon. Not. R. Astron. Soc., 374, 372-384 (2007) New brown dwarfs in Upper Sco using UKIDSS Galactic Cluster Survey science verification data.

J/MNRAS/374/399/ tablea1.dat tableb1.dat (95Kb) Froebrich D., Scholz A., Raftery C.L. Mon. Not. R. Astron. Soc., 374, 399-408 (2007) A systematic survey for infrared star clusters with —b— 20^{∘}using 2MASS.

J/MNRAS/374/664/ tablea1.dat tablea2.dat (72Kb) Cenarro A.J., Peletier R.F., Sanchez-Blazquez P., Selam S.O., Toloba E., Cardiel N., Falcon-Barroso J., Gorgas J., Jimenez-Vicente J., Vazdekis A. Mon. Not. R. Astron. Soc., 374, 664-690 (2007) Medium-resolution Isaac Newton Telescope library of empirical spectra - II. The stellar atmospheric parameters.

J/MNRAS/375/115/ table1.dat table2.dat table3.dat (78Kb) Hwang H.S., Serjeant S., Lee M.G., Lee K.H., White G.J. Mon. Not. R. Astron. Soc., 375, 115-127 (2007) The ultraluminous and hyperluminous infrared galaxies in the Sloan Digital Sky Survey, 2dF Galaxy Redshift Survey and 6dF Galaxy Survey.

J/MNRAS/375/145/ table.dat (27Kb)

Naze Y., Rauw G., Pollock A.M.T., Walborn N.R., Howarth D. Mon. Not. R. Astron. Soc. 375, 145 (2007) Towards an understanding of the Of?p star HD 191612: Phase-resolved multiwavelength observations.

J/MNRAS/375/931/ table3.dat (793Kb) Mauch T., Sadler E.M. Mon. Not. R. Astron. Soc., 375, 931-950 (2007) Radio sources in the 6dFGS: local luminosity functions at 1.4GHz for star-forming galaxies and radio-loud AGN.

J/MNRAS/375/989/ table1.dat (17Kb) Wold M., Brotherton M.S., Shang Z. Mon. Not. R. Astron. Soc., 375, 989-999 (2007) The dependence of quasar variability on black hole mass.

J/MNRAS/375/1220/ cepob3b.dat ic348.dat ngc2264.dat ngc2547.dat ngc7160.dat persei.dat samples.dat sigori.dat (38.1 Mb) Mayne N.J., Naylor T., Littlefair S.P., Saunders E.S., Jeffries R.D. Mon. Not. R. Astron. Soc., 375, 1220-1240 (2007) Empirical isochrones and relative ages for young stars, and the radiative-convective gap.

J/MNRAS/375/1349/ table1.dat (10Kb) Cruz M.J., Jarvis M.J., Rawlings S., Blundell K.M. Mon. Not. R. Astron. Soc., 375, 1349-1363 (2007) The 6C** sample of steep-spectrum radio sources - II. Redshift distribution and the space density of highredshift radio galaxies.

J/MNRAS/376/125/ table1.dat table2.dat table3.dat table4.dat (22Kb) Sanchez S.F., Cardiel N., Verheijen M.A.W., Pedraz S., Covone G. Mon. Not. R. Astron. Soc., 376, 125-150 (2007) Morphologies and stellar populations of galaxies in the core of Abell 2218.

J/MNRAS/376/313/ table1.dat table2.dat (17Kb) Groenewegen M.A.T., Wood P.R., Sloan G.C., Blommaert J.A.D.L., Cioni M.-R.L., Feast M.W., Hony S., Matsuura M., Menzies J.W., Olivier E.A., Vanhollebeke E., Van Loon J.TH., Whitelock P.A., Zijlstra A.A., Habing H.J., Lagadec E. Mon. Not. R. Astron. Soc., 376, 313-337 (2007) Luminosities and mass-loss rates of carbon stars in the Magellanic Clouds.

J/MNRAS/376/361/ table3.dat table4.dat table5.dat table6.dat table7.dat (185Kb) Folsom C.P., Wade G.A., Bagnulo S., Landstreet J.D. Mon. Not. R. Astron. Soc., 376, 361-370 (2007) Rotation and chemical abundances of Ap/Bp stars in the open cluster NGC 6475.

J/MNRAS/376/371/ table2.dat (1.0 Mb) Jackson N., Battye R.A., Browne I.W.A., Joshi S., Muxlow T.W.B., Wilkinson P.N. Mon. Not. R. Astron. Soc., 376, 371-377 (2007) A survey of polarization in the JVAS/CLASS flat-spectrum radio source surveys - I. The data and catalogue production.

J/MNRAS/376/573/ refs.dat table4.dat (13Kb) Ibanoglu C., Tas G., Sipahi E., Evren S. Mon. Not. R. Astron. Soc., 376, 573-579 (2007) VZ CVn: a close binary system with γDoradus-like variations.

J/MNRAS/376/960/ table2.dat table3.dat table4.dat (24Kb) Li Z., Wang Q.D., Hameed S. Mon. Not. R. Astron. Soc., 376, 960-976 (2007) Chandra and XMM-Newton detection of large-scale diffuse X-ray emission from the Sombrero galaxy.

J/MNRAS/376/1073/ cluster.dat tablea1.dat tableb1.dat (31Kb) Zemcov M., Borys C., Halpern M., Mauskopf P., Scott D. Mon. Not. R. Astron. Soc., 376, 1073-1098 (2007) A study of the Sunyaev-Zel’dovich increment using archival SCUBA data.

J/MNRAS/376/1251/ gmrtfls.dat (344Kb) Garn T., Green D.A., Hales S.E.G., Riley J.M., Alexander P. Mon. Not. R. Astron. Soc., 376, 1251-1260 (2007) Deep 610-MHz Giant Metrewave Radio Telescope observations of the Spitzer extragalactic First Look Survey field. I. Observations, data analysis and source catalogue.

J/MNRAS/376/1707/ table1.dat (45Kb) Qiao R.C., Yan Y.R., Shen K.X., Dourneau G., Xi X.J., Cheng X., Wang S.H., Tang Z.H., Liu J.R. Mon. Not. R. Astron. Soc., 376, 1707-1710 (2007) Astrometric observations of Triton.

J/MNRAS/377/430/ table2.dat table3.dat (26Kb) Mckean J.P., Browne I.W.A., Jackson N.J., Fassnacht C.D., Helbig P. Mon. Not. R. Astron. Soc., 377, 430-440 (2007) The Cosmic Lens All-Sky Survey parent population. I. Sample selection and number counts.

J/MNRAS/377/571/ table3.dat (16Kb) Ellingsen S.P. Mon. Not. R. Astron. Soc., 377, 571-583 (2007) A GLIMPSE-based search for 6.7-GHz methanol masers and the lifetime of their spectral features.

J/MNRAS/377/741/ table1.dat (32Kb) Irwin J., Hodgkin S., Aigrain S., Hebb L., Bouvier J., Clarke C., Moraux E., Bramich D.M. Mon. Not. R. Astron. Soc., 377, 741-758 (2007) The Monitor project: rotation of low-mass stars in the open cluster NGC 2516.

J/MNRAS/377/787/ lrgs.dat (4.8 Mb) Barber T., Meiksin A., Murphy T. Mon. Not. R. Astron. Soc. 377, 787 (2007) Properties of Luminous Red Galaxies in the Sloan Digital Sky Survey.

J/MNRAS/377/1737/ table1.dat (150Kb) Piatti A.E., Claria J.J., Mermilliod J.-C., Parisi M.C., Ahumada A.V. Mon. Not. R. Astron. Soc., 377, 1737-1746 (2007) CCD BVI photometry and Coravel observations of stars in the open cluster NGC 2489.

J/MNRAS/378/179/ refs.dat table1.dat table2.dat table3.dat table4.dat table5.dat (104Kb) Bulut I., Demircan O. Mon. Not. R. Astron. Soc. 378, 179 (2007) A New Catalog of Eclipsing Binary Stars with Eccentric Orbits.

J/MNRAS/378/318/ table6.dat table7.dat table8.dat table9.dat (26Kb) Bosler T.L., Smecker-Hane T.A., Stetson P.B. Mon. Not. R. Astron. Soc., 378, 318-338 (2007) The chemical abundances of the stellar populations in the Leo I and II dSph galaxies.

J/MNRAS/378/995/ table1.dat (20Kb) Moss D., Seymour N., McHardy I.M., Dwelly T., Page M.J., Loaring N.S. Mon. Not. R. Astron. Soc., 378, 995-1006 (2007) A deep Giant Metre-wave Radio Telescope 610-MHz survey of the 1^{H} XMM-Newton/Chandra survey field.

J/MNRAS/378/1141/ table1.dat table2.dat table3.dat table4.dat table5.dat table8.dat (40Kb) Gonzalez G., Laws C. Mon. Not. R. Astron. Soc., 378, 1141-1152 (2007) Parent stars of extrasolar planets. VIII. Chemical abundances for 18 elements in 31 stars.

J/MNRAS/379/260/ table1.dat (13Kb) Magliocchetti M., Bruggen M. Mon. Not. R. Astron. Soc., 379, 260-274 (2007) The interplay between radio galaxies and cluster environment.

J/MNRAS/379/275/ table2.dat (7Kb) Perez-Torres M.A., Alberdi A. Mon. Not. R. Astron. Soc., 379, 275-281 (2007) Radio emission from the Sy 1.5 galaxy NGC 5033.

J/MNRAS/379/289/ table1.dat table2.dat table3.dat (39Kb) Kerton C.R., Murphy J., Patterson J. Mon. Not. R. Astron. Soc., 379, 289-296 (2007) A sharper view of the outer Galaxy at 1420 and 408 MHz from the Canadian Galactic Plane Survey. II: the catalogue of extended radio sources.

J/MNRAS/379/647/ table3.dat table4.dat (24Kb) Lister T.A., West R.G., Wilson D.M., Cameron A.C., Clarkson W.I., Street R.A., Enoch B., Parley N.R., Christian D.J., Kane S.R., Evans A., Fitzsimmons A., Haswell C.A., Hellier C., Hodgkin S.T., Horne K., Irwin J., Keenan F.P., Norton A.J., Osborne J., Pollacco D.L., Ryans R., Skillen I., Wheatley P.J., Barnes J.R. Mon. Not. R. Astron. Soc., 379, 647-662 (2007)

SuperWASP-North extrasolar planet candidates: candidates from fields 17h RA 18h.

J/MNRAS/379/663/ table1.dat (20Kb) Moore T.J.T., Bretherton D.E., Fujiyoshi T., Ridge N.A., Allsopp J., Hoare M.G., Lumsden S.L., Richer J.S. Mon. Not. R. Astron. Soc., 379, 663-673 (2007) The star-forming content of the W3 giant molecular cloud.

J/MNRAS/379/816/ table2.dat table3.dat (25Kb) Street R.A., Christian D.J., Clarkson W.I., Cameron A.C., Enoch B., Kane S.R., Lister T.A., West R.G., Wilson D.M., Evans A., Fitzsimmons A., Haswell C.A., Hellier C., Hodgkin S.T., Horne K., Irwin J., Keenan F.P., Norton A.J., Osborne J., Pollacco D.L., Ryans R., Skillen I., Wheatley P.J., Barnes J. Mon. Not. R. Astron. Soc., 379, 816-832 (2007) SuperWASP-N extrasolar planet candidates between 18 RA 21h.

J/MNRAS/379/1237/ table1.dat (8Kb) Maheswar G., Sharma S., Biman J.M., Pandey A.K., Bhatt H.C. Mon. Not. R. Astron. Soc., 379, 1237-1247 (2007) Champagne flow and triggered star formation in NGC 1893.

J/MNRAS/379/1546/ table2.dat (25Kb) Tanaka M., Hoshi T., Kodama T., Kashikawa N. Mon. Not. R. Astron. Soc., 379, 1546-1556 (2007) A huge filamentary structure at z = 0.55 and star formation histories of galaxies at z 1.

J/MNRAS/380/1098/ age.dat lambda.dat sed.dat table2.dat (4.5 Mb) Han Z., Podsiadlowski P., Lynas-Gray A.E. Mon. Not. R. Astron. Soc., 380, 1098-1118 (2007) A binary model for the UV-upturn of elliptical galaxies.

J/MNRAS/381/1219/ fig4 1.fit fig4 2.fit fig4 3.fit table1.dat table2.dat table3.dat (6.2 Mb) Ojha D.K., Tej A., Schultheis M., Omont A., Schuller F. Mon. Not. R. Astron. Soc. 381, 1219 (2007) High mass-loss asymptotic giant branch stars detected by the Midcourse Space Experiment in the ”intermediate” and ”outer” Galactic bulge.

J/MNRAS/382/279/ xlss.dat (767Kb) Pierre M., Chiappetti L., Pacaud F., Gueguen A., Libbrecht C., Altieri B., Aussel H., Gandhi P., Garcet O., Gosset E., Paioro L., Ponman T.J., Read A.M., Refregier A., Starck J.-L., Surdej J., Valtchanov I., Adami C., Alloin D., Alshino A., Andreon S., Birkinshaw M., Bremer M., Detal A., Duc P.-A., Galaz G., Jones L., Le Fevre J.-P., Le Fevre O., Maccagni D., Mazure A., Quintana H., Rottgering H.J.A., Sprimont P.-G., Tasse C., Trinchieri G., Willis J.P. Mon. Not. R. Astron. Soc. 382, 279 (2007) The XMM-LSS catalogue: X-ray sources and associated optical data. Version I.

J/MNRAS/382/382 See VIII/82/

J/MNRAS/382/412/ catalog.dat (2.9 Mb) Ofek E.O., Oguri M., Jackson N., Inada N., Kayo I. Mon. Not. R. Astron. Soc. 382, 412 (2007) SDSS J131339.98+515128.3: A new gravitationally lensed quasar selected on near-infrared excess.

J/PASJ/33/107/  notes.dat table.dat  (67Kb)

Ichikawa T. Publ. Astron. Soc. Jap., 33, 107 (1981) The space distribution of M giants in the direction of Cas.

J/PASJ/51/791/  refs.dat table1.dat table2.dat table3.dat table4.dat  (30Kb)

Yamaguchi R., Saito H., Mizuno N., Mine Y., Mizuno A., Ogawa H., Fukui Y. Publ. Astron. Soc. Jap., 51, 791 (1999) Molecular clouds and star formation in the southern HII regions

J/PASJ/52/997/  table1.dat  (39Kb)

Hirabayashi H., Fomalont E., Horiuchi S., Lovell E.J., Moellenbrock G.A., Inoue M., Burke B.F., Dewdney P.E., Gurvits L.I., Kobayashi H., Jauncey D.L., Murata Y., McCulloch P., Preston R.A., Avruch I.M., Edwards P.G., Dougherty S.M., Scott W.K., Frey S., Paragi Z., Kovalev Y.A., Popov M., Romney J.D., Schilizzi R.T., Shen Z.-Q., Nicolson G., Quick J., Costa M., Dodson R., Reynolds J.E., Tzioumis A.K., Tingay S.J., Hong X.-Y., Liang S.-G., Huang X.-Y., Wei W.-R., Trigilio C., Tuccari G., Nakajima J., Kawai E., Umemoto T., Miyaji T., Fujisawa K., Kawaguchi N., Kus A., Ghigo F., Salter C., Ghosh T., Kanevsky B., Slysh V., Gunn A., Burgess P., Carlson B., del Rizzo D., Taylor R., Cannon W., Kameno S., Shibata K.M., Benson J., Flatters C., Hale A., Lewis C., Langston G., Minter A., Miller K., Smith J., Wietfeldt R., Altunin V., Meier D.L., Murphy D.W., Resch G., Lister M.L., Piner B.G., Jenkins R., Border J., Gimeno J. Publ. Astron. Soc. Jap., 52, 997 (2000) The VSOP 5 GHz AGN survey. I. Compilation and observations.

J/PASJ/53/293/  table1.dat table2.dat  (14Kb)

Deguchi S., Fujii T., Matsumoto S., Nakashima J., Wood P.R. Publ. Astron. Soc. Jap., 53, 293 (2001) Near-IR observations of IRAS sources with SiO masers at the inner Galactic Bulge

J/PASJ/53/445/  table2.dat  (12Kb)

Ishisaki Y., Ueda Y., Yamashita A., Ohashi T., Lehmann I., Hasinger G. Publ. Astron. Soc. Jap., 53, 445 (2001) ASCA Deep survey in the Lockman Hole Field

J/PASJ/53/517/  refs.dat table1.dat table2.dat table3.dat  (49Kb)

Takaba H., Iwata T., Miyaji T., Deguchi S. Publ. Astron. Soc. Jap., 53, 517 (2001) Kashima 34-m water maser survey of late-type stars

J/PASJ/53/653/  table3.dat  (18Kb)

Tamura N., Ohta K., Maihara T., Iwamuro F., Motohara K., Takata T., Iye M. Publ. Astron. Soc. Jap., 53, 653 (2001) NIR narrow- and broad-band study of the SSA 22 field.

J/PASJ/53/971/  table1.dat table2.dat table3.dat  (16Kb)

Mizuno N., Yamaguchi R., Mizuno A., Rubio M., Abe R., Saito H., Onishi T., Yonekura Y., Yamaguchi N., Ogawa H., Fukui Y. Publ. Astron. Soc. Jap., 53, 971 (2001) A CO survey of the LMC with NANTEN: II. Catalog of molecular clouds.

J/PASJ/53/1081/  table1.dat table2.dat  (20Kb)

Tachihara K., Toyoda S., Onishi T., Mizuno A., Fukui Y., Neuhaeuser R. Publ. Astron. Soc. Jap., 53, 1081 (2001) 12CO molecular cloud survey and global star formation in Lupus

J/PASJ/54/719/  notes.dat table1.dat  (21Kb)

Deguchi S., Fujii T., Nakashima J., Wood P.R. Publ. Astron. Soc. Jap., 54, 719 (2002) Near-infrared observations of the IRAS/SiO sources in the Galactic bulge: a large scale distribution.

J/PASJ/55/161/  table1.dat table5.dat table6.dat table8.dat  (51Kb)

Yokogawa J., Imanishi K., Tsujimoto M., Koyama K., Nishiuchi M. Publ. Astron. Soc. Jap., 55, 161 (2003) Review of discrete X-ray sources in the Small Magellanic Cloud : summary of the ASCA results and implication on the recent star-forming activity.

J/PASJ/55/203/  refs.dat table1.dat table2.dat table4.dat table5.dat  (45Kb)

Nakashima J., Deguchi S. Publ. Astron. Soc. Jap., 55, 203 (2003) SiO maser survey toward the inner Galactic disk: 40^{∘}l 70^{∘}and —b— 10^{∘}.

J/PASJ/55/229/  refs.dat table1.dat table2.dat table3.dat table5.dat table6.dat  (37Kb)

Nakashima J., Deguchi S. Publ. Astron. Soc. Jap., 55, 229 (2003) SiO maser survey of cold IRAS sources.

J/PASJ/55/351/  table2.dat table3.dat  (27Kb)

Tingay S.J., Jauncey D.L., King E.A., Tzioumis A.K., Lovell J.E.J., Edwards P.G. Publ. Astron. Soc. Jap., 55, 351-384 (2003) ATCA monitoring observations of 202 compact radio sources in support of the VSOP AGN survey.

J/PASJ/55/635/  table1.dat  (28Kb)

Nakajima H., Imanishi K., Takagi S., Koyama K., Tsujimoto M. Publ. Astron. Soc. Jap., 55, 635 (2003) X-ray observation on the Monoceros R2 star-forming region with the Chandra ACIS-I array.

J/PASJ/55/653/  table2.dat table3.dat  (67Kb)

Imanishi K., Nakajima H., Tsujimoto M., Koyama K., Tsuboi Y. Publ. Astron. Soc. Jap., 55, 653 (2003) A systematic study of X-ray flares from low-mass young stellar objects in the rho Ophiuchi star-forming region

with Chandra.

J/PASJ/55/947/  position.dat table1.dat  (17Kb)

Liang E.-W. Publ. Astron. Soc. Jap., 55, 947-951 (2003) A possible signature of connection between blazars and Seyfert galaxies.

J/PASJ/55/1079/  table4.dat  (28Kb)

Miyazaki M., Shimasaku K., Kodama T., Okamura S., Furusawa H., Ouchi M., Nakata F., Doi M., Hamabe M., Kimura M., Komiyama Y., Miyazaki S., Nagashima C., Nagata T., Nagayama T., Nakajima Y., Nakaya H., Pickles A.J., Sato S., Sekiguchi K., Sekiguchi M., Sugitani K., Takata T., Tamura M., Yagi M., Yasuda N. Publ. Astron. Soc. Jap., 55, 1079-1103 (2003) Evolution of elliptical galaxies at z =1 revealed from a large, multicolor sample of extremely red objects.

J/PASJ/56/261/  table2.dat table3.dat  (33Kb)

Deguchi S., Imai H., Fujii T., Glass S., Ita Y., Izumiura H., Kameya O., Miyazaki A., Nakada Y., Nakashima J. Publ. Astron. Soc. Jap., 56, 261-294 (2004) SiO maser survey of the large-amplitude variables in the Galactic center.

J/PASJ/56/295/  clusters.dat table3.dat  (135Kb)

Hasegawa T., Malasan H.L., Kawakita H., Obayashi H., Kurabayashi T., Nakai T., Hyakkai M., Arimoto N. Publ. Astron. Soc. Jap., 56, 295-311 (2004) New photometric data of old open clusters in the anti-galactic center region.

J/PASJ/56/313/  groups.dat table1.dat table2.dat  (36Kb)

Kim B.G., Kawamura A., Yonekura Y., Fukui Y. Publ. Astron. Soc. Jap., 56, 313-339 (2004) 13CO (J=1-0) survey toward of molecular clouds toward the Monoceros and Canis Major Region.

J/PASJ/56/765/  T1 T3 refs.dat table1.dat table2.dat table3.dat  (64Kb)

Degychi S., Fujii T., Glass I.S., Imai H., Ita Y., Izumiura H., Kameya O., Miyazaki A., Nakada Y., Nakashima J.-I. Publ. Astron. Soc. Jap., 56, 765-802 (2004) SiO maser survey of IRAS sources in the inner galactic disk.

J/PASJ/56/965/  table3.dat table4.dat table5.dat table6.dat table7.dat  (150Kb)

Fukazawa Y., Makishima K., Ohashi T. Publ. Astron. Soc. Jap., 56, 965-1009 (2004) ASCA compilation of X-ray properties of hot gas in elliptical galaxies and galaxy clusters: two breaks in the temperature dependences.

J/PASJ/56/1025/  codes.dat table1.dat table2.dat  (19Kb)

Kodaira K., Vansevicius V., Bridzius A., Komiyama Y., Miyazaki S., Stonkute R., Sableviciute I., Narbutis D. Publ. Astron. Soc. Jap., 56, 1025-1040 (2004) A survey of compact star clusters in the south-west field of the M 31 disk.

J/PASJ/57/13/  raw/* reduc/* table1.dat table2.dat  (264.4 Mb)

Takeda Y., Sato B., Kambe E., Masuda S., Izumiura H., Watanabe E., Ohkubo M., Yanagisawa K., Yasaka Y., Honda S., Kawanomoto S., Aoki W., Ando H., Sadakane K., Takada-hidai M., Miyama S.M. Publ. Astron. Soc. Jap., 57, 13-25 (2005) High-dispersion spectra collection of nearby F-K stars at Okayama astrophysical observatory: a basis for spectroscopic abundance standards.

J/PASJ/57/27/  etab2/* table1.dat tablee1.dat tablee2.dat tablee2a.dat tablee3.dat  (3.5 Mb)

Takeda Y., Ohkubo M., Sato B., Kambe E., Sadakane K. Publ. Astron. Soc. Jap., 57, 27-43 (2005) Spectroscopic study on the atmospheric parameters of nearby F-K dwarfs and subgiants.

J/PASJ/57/45/  ncor m05.dat ncor m10.dat ncor p00.dat ncor p05.dat table1.dat  (82Kb)

Takeda Y., Kawanomoto S. Publ. Astron. Soc. Jap., 57, 45-63 (2005) Lithium abundances of F-, G-, and K-type stars: profile-fitting analysis of the Li I 6708 doublet.

J/PASJ/57/65/ ackerman.dat c9061vt1.dat c9061vt2.dat c9061vt3.dat c9078vt1.dat c9078vt2.dat c9078vt3.dat c9088vt1.dat c9088vt2.dat c9088vt3.dat c9094vt1.dat c9094vt2.dat c9094vt3.dat c9111vt1.dat c9111vt2.dat c9111vt3.dat cabund.dat fitabund.dat nabund.dat oabund.dat tomkin92.dat tomkin95.dat (623Kb) Takeda Y., Honda S. Publ. Astron. Soc. Jap., 57, 65-82 (2005)

Photospheric CNO abundances of solar-type stars.

J/PASJ/57/109/ average.dat h104985.dat h104985a.dat table1.dat table2.dat tablee1.dat tablee1a.dat tablee2.dat (1.1 Mb) Takeda Y., Sato B., Kambe E., Izumiura H., Masuda S., Ando H. Publ. Astron. Soc. Jap., 57, 109-125 (2005) Stellar parameters and photospheric abundances of late-G giants: properties of the targets of the Okayama planet search program.

J/PASJ/57/135/  table8.dat  (15Kb)

Inui T., Matsumoto H., Tsuru T.G., Koyama K., Matsushita S., Peck A.B., Tarchi A. Publ. Astron. Soc. Jap., 57, 135-145 (2005) Chandra observation of the starburst galaxy NGC 2146.

J/PASJ/57/165/  table2.dat  (10Kb)

Taniguchi Y., Ajiki M., Nagao T., Shioya Y., Murayama T., Kashikawa N., Kodaira K., Kaifu N., Ando H., Karoji H., Akiyama M., Aoki K., Doi M., Fujita S.S., Furusawa H., Hayashino T., Iwamuro F., Iye M., Kobayashi N., Kodama T., Komiyama Y., Matsuda Y., Miyazaki S., Mizumoto Y., Morokuma T., Motohara K., Nariai K., Ohta K., Ohyama Y., Okamura S., Ouchi M., Sasaki T., Sato Y., Sekiguchi K., Shimasaku K., Tamura H., Umemura M., Yamada T., Yasuda N., Yoshida M. Publ. Astron. Soc. Jap., 57, 165-182 (2005) The SUBARU deep field project: Lyman-αemitters at a redshift of 6.6.

J/PASJ/57/881/  table1.dat table3.dat table4.dat  (42Kb)

Yamada S.F., Sasaki S.S., Sumiya R., Umeda K., Shioya Y., Ajiki M., Nagao T., Murayama T., Taniguchi Y. Publ. Astron. Soc. Jap., 57, 881-903 (2005) An intermediate-band imaging survey for high-redshift Lyman alpha emitters: the Mahoroba-11.

J/PASJ/57/917/  table1.dat  (11Kb)

Nakagawa M., Onishi T., Mizuno A., Fukui Y. Publ. Astron. Soc. Jap., 57, 917-931 (2005) An unbiased search for molecular clouds in the southern Galactic Warp.

J/PASJ/58/L29/  tablee1.dat  (10Kb)

Matsuyanagi I., Itoh Y., Sugitani K., Oasa Y., Mukai T., Tamura M. Publ. Astron. Soc. Jap., 58, L29-L34 (2006) Sequential formation of low-mass stars in the BRC 14 region.

J/PASJ/58/313/  table1.dat table3.dat  (18Kb)

Shimasaku K., Kashikawa N., Doi M., Ly C., Malkan M.A., Matsuda Y., Ouchi M., Hayashino T., Iye M., Motohara K., Murayama T., Nagao T., Ohta K., Okamura S., Sasaki T., Shioya Y., Taniguchi Y. Publ. Astron. Soc. Jap., 58, 313-334 (2006) Ly-alpha emitters at z = 5.7 in the Subaru deep field.

J/PASJ/58/389/ dhpeg-1.dat dhpeg-2.dat dhpeg-3.dat dxdel-1.dat dxdel-2.dat dxdel-3.dat ewidlist.dat rrlyr-1.dat rrlyr-2.dat rrlyr-3.dat rrlyr-4.dat rrlyr-5.dat solarfe.dat vyser-1.dat vyser-2.dat vyser-3.dat vyser-4.dat (142Kb) Takeda Y., Honda S., Aoki W., Takada-hidai M., Zhao G., Chen Y.-Q., Shi J.-R. Publ. Astron. Soc. Jap., 58, 389-406 (2006) On the spectroscopic determination of atmospheric parameters and O/Fe abundances of RR Lyrae stars.

J/PASJ/58/529/  table1.dat table3.dat  (46Kb)

Fujii T., Deguchi S., Ita Y., Izumiura H., Kameya O., Miyazaki A., Nakada Y. Publ. Astron. Soc. Jap. 58, 529 (2006) SiO Maser Survey of the Inner Bar of the Galactic Bulge.

J/PASP/(in/press See II/271A/

J/PASP//2007 See II/271A/

J/PASP/83/611 See VII/11/

J/PASP/85/70 See II/212A/

J/PASP/87/969 See II/63A/

J/PASP/97/753/  table6.dat  (62Kb)

Mateo M., Hodge P. Publ. Astron. Soc. Pac. 97, 753 (1985) CCD photometry of Large Magellanic Cloud clusters. I. The remote cluster NGC 1777.

J/PASP/97/932/  3cr.dat ref.dat  (45Kb)

Spinrad. H, Djorgovski S., Marr J., Aguilar L. Publ. Astron. Soc. Pac., 97, 932 (1985) 3CR Source Identifications, Update 1990

J/PASP/98/192 See II/117/

J/PASP/98/572 See II/220/

J/PASP/98/1133/  table1  (7Kb)

NIEMELA V.S., MARRACO H.G, CABANNE M.L. Publ. Astron. Soc. Pac. 98, 1133 (1986) The young open cluster NGC 346 in the Small Magellanic Cloud

J/PASP/99/695 See VI/42/

J/PASP/99/739/  tables.dat  (369Kb)

Hesser J.E., Harris W.E., Vandenberg D.A., Allwright J.W.B., Shott P., Stetson B. Publ. Astron. Soc. Pac. 99, 739 (1987) A CCD color-magnitude study of 47 Tucanae.

J/PASP/99/1184 See II/138/

J/PASP/102/379 See III/195/

J/PASP/102/773 See II/220/

J/PASP/102/1281 See VII/139/

J/PASP/102/1392/  table2.dat table4.dat  (20Kb)

Doinidis S.P., Beers T.C. Publ. Astron. Soc. Pac. 102, 1392 (1990) Photoelectric UBV photometry of Northern stars from the HK objective-prism survey

J/PASP/103/209 See VII/133/

J/PASP/103/536/  table1  (11Kb)

CRINKLAW G., TALBERT F.D. Publ. Astron. Soc. Pac. 103, 536 (1991) CCD Photometry of the old open cluster NGC 7142

J/PASP/103/973/  table4.dat  (25Kb)

Doidinis S.P., Beers T.C. Publ. Astron. Soc. Pac. 103, 973 (1991) Photoelectric UBV photometry of northern stars from the HK objective-prism survey. II.

J/PASP/104/533 See II/179/

J/PASP/104/553/  adc.sty odewahn.tex tables  (32Kb)

ODEWAHN S.C., BRYJA C., HUMPHREYS R.M. Publ. Astron. Soc. Pac. 104, 553 (1992) Improved CCD Standard Fields

J/PASP/104/678 See VII/149/

J/PASP/104/1063/  adc.sty table4 walker.tex walker.txt  (107Kb)

Walker A.R. Publ. Astron. Soc. Pac. 104, 1063 (1992) A BV Color-Magnitude Diagram for the Galactic Globular Cluster NGC 1851

J/PASP/104/1107 See II/187/

J/PASP/105/36/  table1 table2  (55Kb)

Sowell J.R., Wilson J.W. Publ. Astron. Soc. Pac. 105, 36 (1993) All-sky Stroemgren photometry of speckle binary stars

J/PASP/105/127 See V/79/

J/PASP/105/693 See III/181/

J/PASP/105/1127/  table2  (43Kb)

Gaustad J.E., Van Buren D. Publ. Astron. Soc. Pac. 105, 1127 (1993) The Distribution of Interstellar Dust in the Solar Neighborhood

J/PASP/105/1279/  table2c.dat  (13Kb)

Hyung S., Aller L. H., Feibelman W. A.: 1993 Publ. Astron. Soc. Pac. 105, 1279-1286 (1993) The Spectrum of the Planetary Nebula NGC 6567

J/PASP/105/1465/  table1.dat  (103Kb)

Reed B.C. Publ. Astron. Soc. Pac. 105, 1465 (1993) Distant (r 5 kpc) OB Stars in the Galaxy

J/PASP/106/38 See II/212A/

J/PASP/106/281/  tables.dat  (38Kb)

Daniel S.A., Latham D.W., Mathieu R.D., Twarog B.A. Publ. Astron. Soc. Pac. 106, 281 (1994) A photometric and radial-velocity analysis of the intermediate-age open cluster NGC 752.

J/PASP/106/382 See III/179/

J/PASP/106/404/  table2.dat  (40Kb)

SARAJEDINI A. Publ. Astron. Soc. Pac. 106, 404 (1994) (From AAS CD-ROM series, Volume 2, 1994) CCD photometry of the galactic globular cluster NGC6535 in the B and V passbands

J/PASP/106/452/  table8.dat  (21Kb)

Taylor B.J. Publ. Astron. Soc. Pac. 106, 452 (1994) Analysis of Archival Data for Cool Dwarfs. II. A Catalog of Temperatures

J/PASP/106/486/  table3.dat  (56Kb)

ANTHONY-TWAROG B.J., TWAROG B.A., SHEERAN M. Publ. Astron. Soc. Pac. 106, 486 (1994) (From AAS CD-ROM series, Volume 2, 1994) vbyHbeta CCD photometry of Melotte 66: a disk analog of Omega Centauri?

J/PASP/106/566 See II/179/

J/PASP/106/646/  notes.dat table1.dat table2.dat  (143Kb)

Kirhakos S., Sargent W.L.W., Schneider D.P., Bahcall J.N., Jannuzi B.T., Maoz D., Small T.A. Publ. Astron. Soc. Pac. 106, 646 (1994) The HST quasar absorption-line key project. VIII. CCD imaging of Hubble Space Telescope quasar fields.

J/PASP/106/704 See V/88/

J/PASP/106/745/  table5.dat  (20Kb)

Hyung S., Aller L. H., and Feibelman W. A. Publ. Astron. Soc. Pac. 106, 745, (1994) The Spectrum of the Planetary Nebula, IC 418

J/PASP/106/828/  table4.dat table5.dat  (16Kb)

Walker A.R. Publ. Astron. Soc. Pac. 106, 828 (1994) BVRI photometry of the omega Centauri Hubble Space Telescope calibration field

J/PASP/106/949 See II/220/

J/PASP/106/967/  table6.dat  (22Kb)

Joergensen Inger Publ. Astron. Soc. Pac., 106, 967 (1994) Secondary Photometric Standard Stars for the Thuan-Gunn and Johnson-Kron-Cousins Systems.

J/PASP/106/1120/  table1 table2 table3  (37Kb)

Milone E.F., Wilson W.J.F., Fry D.J.I., Schiller S.J. Publ. Astron. Soc. Pac. 106, 1120 (1994) (From AAS CD-ROM series, Volume 3, 1995) Studies of large amplitude delta Scuti variables II: DY Herculis

J/PASP/106/1151 See VI/104/

J/PASP/107/32/  table3  (54Kb)

Anthony-Twarog B.J., Twarog B.A., Craig J. Publ. Astron. Soc. Pac. 107, 32 (1995) CN and Ca abundance variations among the giants in M 22

J/PASP/107/299/  notes.dat occulcat.dat occulcat.doc refs.dat  (122Kb)

Mason B.D. Publ. Astron. Soc. Pac. 107, 299 (1995) Catalog of Photoelectric Measures of Occultation Binary Stars

J/PASP/107/324/  table2 table3  (43Kb)

Henden A.A., Honeycutt R.K. Publ. Astron. Soc. Pac. 107, 324 (1995) Secondary Photometric Standards for Northern Nova-Like Cataclysmic Variables

J/PASP/107/617/  table1  (48Kb)

Pollanen M.D., Feldman P.A. Publ. Astron. Soc. Pac. 107, 617 (1995) A catalog of the youngest YSOs and candidate protostars.

J/PASP/107/734 See V/88/

J/PASP/107/751 See II/212A/

J/PASP/107/846/  notes.dat table1.dat  (23Kb)

Drilling J.S., Bergeron L.E. Publ. Astron. Soc. Pac. 107, 846 (1995) An extension of the Case-Hamburg OB-star surveys.

J/PASP/107/1042/  ( total of 7 files)  (37.6 Mb)

Hinkle K., Wallace L., Livingston W. Publ. Astron. Soc. Pac. 107, 1042 (1995) (From AAS CD-ROM series, Volume 5, 1995 Lee Brotzman [ADS] 07-Nov-1995) Infrared Atlas of the Arcturus Spectrum, 0.9-5.3 microns

J/PASP/108/277 See III/211/

J/PASP/108/313/  blanket.dat table4.dat table5.dat table6.dat  (47Kb)

Hall J.C. Publ. Astron. Soc. Pac. 108, 313 (1996) (From AAS CD-ROM series, Volume 6, 1996 Lee Brotzman [ADS] 07-May-1996) On the determination of empirical stellar flux scales

J/PASP/108/332/  table1a.dat table1b.dat table1c.dat table3a.dat table3b.dat table3c.dat  (67Kb)

Lapasset E., Gomez M., Farinas R. Publ. Astron. Soc. Pac. 108, 332 (1996) (From AAS CD-ROM series, Volume 6, 1996 Lee Brotzman [ADS] 07-May-1996) Photometric analyses of the short-period contact binaries HY Pavonis, AW Virginis, and BP Velorum

J/PASP/108/338/  table1b.dat table1u.dat table1v.dat  (28Kb)

Gonzalez J.F., Lapasset E., Gomez M., Ahumada J. Publ. Astron. Soc. Pac. 108, 338 (1996) (From AAS CD-ROM series, Volume 6, 1996 Lee Brotzman [ADS] 07-May-1996) Photometric analysis of the contact binary BF Pavonis

J/PASP/108/682/  catalog.dat  (21Kb)

Mould J. R., Watson A. M., Gallagher III J. S., Ballester G. E., Burrows, C. J., Casertano S., Clarke J. T., Crisp D., Griffiths R. E., Hester J., Hoessel J. G., Holtzman J. A., Scowen P. A., Stapelfeld K. R., Trauger J. T., Westphal J. A. Publ. Astron. Soc. Pac., 108, 682, (1996) Low-Mass Stars in Outer Field in NGC 6397

J/PASP/108/1105/  table1  (55Kb)

Schmidt E.G. Publ. Astron. Soc. Pac. 108, 1105 (1996) (From AAS CD-ROM series, Volume 7, 1996 Lee Brotzman [ADS] 09-Nov-1996) Stars classified as constant in the General Catalogue of Variable Stars. II.

J/PASP/109/9/  table2.dat table3.dat table4.dat table5.dat tablea  (44Kb)

Adelman S.J. Publ. Astron. Soc. Pac. 109, 9 (1997) (From AAS CD-ROM series, Volume 7, 1996 Lee Brotzman [ADS] 04-Nov-1996) Stromgrem uvby photometry of the magnetic chemically peculiar stars HR 1643, Theta Aur, 49 Cam, and HR 3724

J/PASP/109/345 See V/94/

J/PASP/109/441/  table3.dat table3.tex  (317Kb)

Henden A.A., Honeycutt R.K. Publ. Astron. Soc. Pac. 109, 441 (1997) (From AAS CD-ROM series, Volume 8, 1997 Lee Brotzman [ADS] 07-May-1997) Secondary Photometric Standards for Northern Cataclysmic Variables and Related Objects

J/PASP/109/524/  table2.dat  (81Kb)

Layden A.C. Publ. Astron. Soc. Pac. 109, 524 (1997) Light Curves For 40 Field RR Lyrae Variables

J/PASP/109/584 See VI/104/

J/PASP/109/614 See III/211/

J/PASP/109/759/  table1.dat table2.dat  (33Kb)

Wolff S.C., Simon T. Publ. Astron. Soc. Pac. 109, 759 (1997) Rotational Velocities of Intermediate-mass Main Sequence Stars

J/PASP/109/849/  table1.dat table2.dat table3.dat table4.dat  (28Kb)

Gizis J.E., Reid, I.N. Publ. Astron. Soc. Pac. 109, 849 (1997) Probing the LHS Catalog I. New Nearby Stars and the Coolest Subdwarf

J/PASP/109/883/  table9.dat table9.ps table9.tex  (394Kb)

Rutledge G.A., Hesser J.E., Stetson P.B., Mateo M., Simard L., Bolte M., Friel E.D., Copin Y. Publ. Astron. Soc. Pac. 109, 883 (1997) (From AAS CD-ROM series, Volume 9, 1997 Lee Brotzman [ADS] 04-Nov-1997) Galactic globular cluster metallicity scale from the Ca II triplet I. Catalog

J/PASP/109/977/  table2.dat  (17Kb)

Layden A.C., Wachter S. Publ. Astron. Soc. Pac. 109, 977 (1997) (From AAS CD-ROM series, Volume 9, 1997 Lee Brotzman [ADS] 04-Nov-1997) CG Muscae: a cataclysmic misclassification

J/PASP/109/998/  table1.dat table2.dat table3.dat  (79Kb)

Shara M.M., Bergeron L.E., Christian C.A., Craig N., Bowyer S. Publ. Astron. Soc. Pac. 109, 998 (1997) An optical atlas of extreme ultraviolet explorer (EUVE) sources

J/PASP/110/533/  table1.dat  (20Kb)

Stetson P.B., Hesser J.E., Smecker-Hane T.A. Publ. Astron. Soc. Pac. 110, 533 (1998) Homogeneous photometry for star clusters and resolved galaxies. I. A survey of bright stars in the Fornax dwarf spheroidal galaxy.

J/PASP/110/553/  table1.dat  (11Kb)

Richmond M.W., Filippenko A.V., Galisky J. Publ. Astron. Soc. Pac. 110, 553 (1998) The supernova rate in starburst galaxies

J/PASP/110/586 See II/212A/

J/PASP/110/779/  position.dat  (300Kb)

Corwin H.G., Pesenson I., Schmitz M., Kim D.-C., Bennett J., Madore B.F., Mazzarella J.M., Helou, G. Publ. Astron. Soc. Pac. 110, 779 (1998) Accurate Positions for MCG Galaxies

J/PASP/110/863/  ( total of 7 files)  (51.7 Mb)

Publ. Astron. Soc. Pac. 110, 863 (1998) A Stellar Spectral Flux Library: 1150 - 25000 A Pickles A.J.

J/PASP/110/1012/  acquis.dat doubles.dat names.dat table2.dat  (8.7 Mb)

Schneider G., Hershey J.L., Wenz M.T. Publ. Astron. Soc. Pac. 110, 1012 (1998) Duplicity in Hubble Space Telescope Guide Stars: Fine Guidance Sensor serendipitous survey results.

J/PASP/110/1172/  table2.dat table3.dat  (33Kb)

Gim M., Hesser J.E., McClure R.D., Stetson P.B. Publ. Astron. Soc. Pac. 110, 1172 (1998) The open cluster NGC 7789. I. Radial velocities for giant stars.

J/PASP/110/1434/  table3.dat  (13Kb)

Stock J., Rose J.A., Agostinho R.J. Publ. Astron. Soc. Pac. 110, 1434 (1998) Digital objective prism spectroscopy and astrometry: quantitative methods and application to high-velocity

stars.

J/PASP/111/57/  table1.dat table3.dat  (29Kb)

Hudson M.J. Publ. Astron. Soc. Pac. 111, 57 (1999) A test for large-scale systematic errors in maps of galactic reddening

J/PASP/111/335/  fig1.gif fig2.gif fig3.gif fig4.gif fig5.gif table1a.dat table1b.dat table1c.dat  (425Kb)

Wallerstein G., Machado-Palaez L., Gonzalez G. Publ. Astron. Soc. Pac. 111, 335 (1999) The Ca II-Mv Correlation (Wilson-Bappu Effect) Calibrated by Hipparcos Parallaxes

J/PASP/111/438/  catalog.dat refs.dat  (1.3 Mb)

Falco E.E., Kurtz M.J., Geller M.J., Huchra J.P., Peters J., Berlind P., Mink D.J., Tokarz S.P., Elwell B. Publ. Astron. Soc. Pac. 111, 438 (1999) The updated Zwicky catalog (UZC).

J/PASP/111/494 See II/220/

J/PASP/111/685/  table1.dat  (59Kb)

Hodge P.W., Balsley J., Wyder T.K., Skelton B.P. Publ. Astron. Soc. Pac. 111, 685 (1999) The HII regions of M33. II. A photometric catalog of 1272 newly identified emission regions.

J/PASP/111/812/  table1.dat  (14Kb)

Barnes T. G. III, Ivans I. I., Martin J. R., Froning C. S., Moffett T. J. Publ. Astron. Soc. Pac., 111, 812 (1999) V(RI)C Photometry of Cepheids in the Magellanic Clouds

J/PASP/111/1169/  apx.dat notes.dat table1.dat  (36Kb)

Westpfahl D.J., Bell D.J., Yoss K.M. Publ. Astron. Soc. Pac. 111, 1169 (1999) A list of late-type stars in the galactic antirotation direction.

J/PASP/111/1233/ table4a.dat table4a.tex table4b.dat table4b.tex table4c.dat table4c.tex table4d.dat table4d.tex (228Kb) Soderberg A.M., Pilachowski C.A., Barden S.C., Willmarth D., Sneden C. Publ. Astron. Soc. Pac. 111, 1233 (1999) Radial Velocities of Giant Stars in Globular Clusters

J/PASP/112/251 See II/231/

J/PASP/112/397 See II/230/

J/PASP/113/10/  table2.dat table3.dat  (21Kb)

Masci F.J., Condon J.J., Barlow T.A., Lonsdale C.J., Xu C., Shupe D.L., Pevunova O., Fang F., Cutri R. Publ. Astron. Soc. Pac. 113, 10 (2001) A new complete sample of submilliJansky radio sources: an optical and near-infrared study.

J/PASP/113/267/  tablea.dat tableb.dat  (75Kb)

Sandage A. Publ. Astron. Soc. Pac. 113, 267 (2001) The Mount Wilson halo mapping project 1975-1985. II. Photometric properties of the Mount Wilson catalogue of photographic magnitudes in Selected Areas 1-139.

J/PASP/113/764 See V/110/

J/PASP/113/1365/  table1.dat  (14Kb)

Buzzoni A., Chavez M., Malagnini M.L., Morossi C. Publ. Astron. Soc. Pac. 113, 1365 (2001) Lick spectral indices for super-metal-rich stars.

J/PASP/113/1393 See II/257/

J/PASP/114/233/  table2.dat  (43Kb)

Uribe A., Garcia-Varela J.-A., Sabogal-Martinez B.-E., Higuera G.M.A., Brieva E. Publ. Astron. Soc. Pac. 114, 233 (2002) Membership in the region of the double cluster h and χPersei working from proper motions and positions: distance moduli and extinction in that galactic direction.

J/PASP/114/298/  table4.dat  (32Kb)

Pietrzynski G., Gieren W., Udalski A. Publ. Astron. Soc. Pac. 114, 298 (2002) Photometric standard stars in the BVI system in a wide field centered on the spiral galaxy NGC 300.

J/PASP/114/656/  table2.dat  (23Kb)

Everett M.E., Howell S.B., van Belle G.T., Ciardi D.R. Publ. Astron. Soc. Pac. 114, 656 (2002) Stellar variability in a survey of field stars.

J/PASP/114/820/  position.dat table1.dat  (20Kb)

van den Bergh S., Li W., Filippenko A.V. Publ. Astron. Soc. Pac. 114, 820 (2002) Classifications of the host galaxies of supernovae.

J/PASP/114/1382/  table1.dat table2.dat table3.dat  (21Kb)

Osborn W., Sano Y., Spalding R. Publ. Astron. Soc. Pac. 114, 1382 (2002) + Erratum 115, 761 (2003) The neglected open cluster Stock 1.

J/PASP/115/22/  table1.dat  (22Kb)

Yong D., Lambert D.L. Publ. Astron. Soc. Pac. 115, 22 (2003) A search for cool subdwarfs: stellar parameters for 134 candidates.

J/PASP/115/351/  table1.dat  (62Kb)

MacConnell D.J. Publ. Astron. Soc. Pac. 115, 351 (2003) Southern Cool Carbon Stars found on near-infrared objective prism plates.

J/PASP/115/413/  catalog.dat crossid.dat  (2.9 Mb)

Stetson P.B., Bruntt H., Grundahl F. Publ. Astron. Soc. Pac., 115, 413-447 (2003) Homogeneous photometry. III. A star catalog for the open cluster NGC 6791.

J/PASP/115/837 See II/254/

J/PASP/115/1280/  table1.dat  (19Kb)

van den Bergh S., Li W., Filippenko A.V. Publ. Astron. Soc. Pac. 115, 1280 (2003) Classifications of the host galaxies of supernovae, set II.

J/PASP/116/622/  table1.dat  (47Kb)

Prochaska J.X., Herbert-Fort S. Publ. Astron. Soc. Pac., 116, 622-633 (2004) The Sloan Digital Sky Survey damped Lyαsurvey: Data Release 1.

J/PASP/116/1012/  cat1.dat cat2.dat crossid.dat  (2.4 Mb)

Stetson P.B., McClure R.D., Vandenberg D.A. Publ. Astron. Soc. Pac., 116, 1012-1030 (2004) A star catalog for the open cluster NGC 188.

J/PASP/116/1039 See II/257/

J/PASP/117/1/  refs.dat table1.dat table2.dat table3.dat  (37Kb)

Dawson P.C., De Robertis M.M. Publ. Astron. Soc. Pac., 117, 1-12 (2005) High proper motion stars. IV. Radial velocities of 166 Luyten Half-Second stars.

J/PASP/117/22/  table1.dat  (660Kb)

Piatti A.E., Claria J.J., Ahumada A.V. Publ. Astron. Soc. Pac., 117, 22-31 (2005) A CCD BVI photometric study of the young, highly reddened open cluster NGC 6318.

J/PASP/117/37/  table4.dat table7.dat table10.dat  (872Kb)

Saha A., Dolphin A.E., Thim F., Whitmore B. Publ. Astron. Soc. Pac., 117, 37-78 (2005) Faint BVRI photometric sequences in selected fields.

J/PASP/117/199/  table1.dat  (15Kb)

Le Bertre T., Tanaka M., Yamamura I., Murakami H., MacConnell D.J. Publ. Astron. Soc. Pac., 117, 199-206 (2005) Carbon stars in the Infrared Telescope in Space Survey.

J/PASP/117/227/  table1.dat table2.dat  (22Kb)

Kewley L.J., Jansen R.A., Geller M.J. Publ. Astron. Soc. Pac., 117, 227-244 (2005) Aperture effects on star formation rate, metallicity, and reddening.

J/PASP/117/676/  table2.dat table4.dat table6.dat  (39Kb)

Reid I.N., Gizis J.E. Publ. Astron. Soc. Pac., 117, 676-698 (2005) Probing the LHS catalog. II. Faint proper-motion stars.

J/PASP/117/773/  table1.dat  (18Kb)

van den Bergh S., Li W., Filippenko A.V. Publ. Astron. Soc. Pac., 117, 773-782 (2005) Classifications of the host galaxies of supernovae, set III.

J/PASP/117/1325/  table7.dat table23.dat table89.dat  (22Kb)

Stetson P.B., Catelan M., Smith H.A. Publ. Astron. Soc. Pac., 117, 1325-1361 (2005) Homogeneous photometry. V. The globular cluster NGC 4147.

J/PASP/118/636/  atlas.dat lines.dat linesne0.dat liness1.dat liness10.dat synt.dat  (19.0 Mb)

Lee B.-C., Galazutdinov G., Han I., Kim K.-M., Yushchenko A.V., Kim J., Tsymbal V., Park M.-G. Publ. Astron. Soc. Pac. 118, 636 (2006) A high resolution spectral atlas of Alpha Persei 3810-8100 Angstroms.

J/PASP/118/1666 See II/271A/

J/PAZh/20/693/  tables.dat  (18Kb)

Evstigneeva N.M., Samus N.N., Tsvetkova T.M., Shokin Yu.A. Pis’ma Astron. Zh. 20, 693 (1994) Identifications and coordinates of variable stars in the globular cluster M 3 = NGC 5272.

J/PAZh/21/13/  table  (10Kb)

Melnik A.M., Efremov Yu.N. Pis’ma Astron. Zh. 21, 13 (1995) New list of OB associations of our Galaxy.

J/PAZh/21/509/  table2.dat  (10Kb)

Evstigneeva N.M., Shokin Yu.A., Samus N.N., Tsvetkova T.M. Pis’ma Astron. Zh. 21, 509 (1995) Identifications and coordinates of variable stars in the globular cluster M 5 = NGC 5904.

J/PAZh/21/596/  table2.dat table3.dat  (18Kb)

Samus’ N.N., Ipatov A.P., Smirnov O.M., Kravtsov V.V., Alcaino G., Liller W., Alvarado F. Pis’ma Astron. Zh., 21, 596 (1995) Photographic photometry of stars and positions of variables in the globular cluster NGC 6584.

J/PAZh/21/652/  table.dat  (18Kb)

Lyuty V.M., Doroshenko V.T., Metlov V.G., Irsmambetova T.R. Pis’ma Astron. Zh. 21, 652 (1995) Rapid optical variability in NGC 7469: UBV photometry.

J/PAZh/22/269/  table2.dat table34.dat  (22Kb)

Samus N.N., Kravtsov V.V., Pavlov M.V., Strel’nikov V.V., Shokin Yu.A., Alcaino G., Liller W., Alvarado F. Pis’ma Astron. Zh. 22, 269 (1996) The giant branch and variable stars in the globular cluster NGC 3201.

J/PAZh/22/764/  table2.dat  (19Kb)

Samus N.N., Kravtsov V.V., Pavlov M.V., Alcaino G., Liller W., Alvarado F. Pis’ma Astron. Zh. 22, 764 (1996) Photographic photometry of stars in the poorly studied globular cluster NGC 6139

J/PAZh/22/847/  table12.dat  (6Kb)

Shokin Y.A., Samus N.N. Pis’ma Astron. Zh. 22, 847 (1996) Identification and coordinates of variable stars in the globular cluster M 4 = NGC 6121.

J/PAZh/23/90/  table.dat  (15Kb)

Glushkova E.V., Zabolotskikh M.V., Rastorguev A.S., Uglova I.M., Fedorova A.A. Pis’ma Astron. Zh. 23, 90 (1997) Absolute proper motions in 181 young open clusters

J/PAZh/23/113/  table1.dat  (6Kb)

Sharov A.S., Lyutyi V.M. Pis’ma Astron. Zh. 23, 113 (1997) Photoelectric observations of Pleione in 1991-1996.

J/PAZh/23/277/  table.dat  (36Kb)

Zaitseva G.V., Lyutyi V.M. Pis’ma Astron. Zh. 23, 277 (1997) On the origin of optical variability of WW Vulpeculae.

J/PAZh/23/341/  table.dat  (12Kb)

Irsmambetova T.R. Pis’ma Astron. Zh. 23, 341 (1997) Analysis of flare activity in SS 433 on the basis of photoelectric photometry.

J/PAZh/23/409/  table1.dat  (16Kb)

Chuvaev K.K., Doroshenko V.T., Sergeev S.G., Pronik V.I, Mal’kov Yu.F. Pis’ma Astron. Zh. 23, 409 (1997) Spectrophotometry of the Seyfert 1 Galaxy Mrk 509 from 1972-1993 near H-beta.

J/PAZh/23/454/  tables.dat  (7Kb)

Evstigneeva N.M., Samus N.N., Alcaino G. Pis’ma Astron. Zh. 23, 454 (1997) Identification and coordinates of variable stars in the globular clusters M 53 = NGC 5024 and NGC 2808.

J/PAZh/23/458/  table1 table2  (13Kb)

Zakirov M.M. Pis’ma Astron. Zh., 23, 458 (1997) Close binary systems in star-forming regions. V502 Cyg in the T-association Cyg T2.

J/PAZh/23/626/  table3a.dat table3b.dat  (22Kb)

Zakirov M.M. Pis’ma Astron. Zh. 23, 626 (1997) Close binary systems in star-forming regions: GG Ori and OS Ori in the Ori I association.

J/PAZh/23/674/  table2.dat  (63Kb)

Glushkova E.V., Uglova I.M. Pis’ma Astron. Zh. 23, 674 (1997) Proper motions of stars in the field of the open cluster NGC 7063

J/PAZh/23/931/  table1.dat table2.dat table3.dat  (126Kb)

Rastorguev A.S., Glushkova E.V. Pis’ma Astron. Zh. 23, 931 (1997) A catalog of radial velocities, positions and proper motions of stars in 19 Kapteyn’s northern areas.

J/PAZh/24/93/  catalog.dat table2.dat  (23Kb)

Shokin Yu.A., Kulagina N.B. Pis’ma Astron. Zh. 24, 93 (1998) A reference catalog for accurate position determinations in the region of M 31

J/PAZh/24/145/  table2.dat  (13Kb)

Zakirov M.M. Pis’ma Astron. Zh. 24, 145 (1998) Close binaries systems in the open cluster regions. V1481 Cyg in NGC 7128.

J/PAZh/24/307/  notes4.dat table1.dat table2.dat table4.dat  (32Kb)

Shatsky N.I. Pis’ma Astron. Zh., 24, 307 (1998) The nature of visual components in 82 multiple systems

J/PAZh/24/343/  table1.dat  (15Kb)

Tokovinin A.A. Pis’ma Astron. Zh., 24, 343 (19980 New spectroscopic components in six multiple systems

J/PAZh/24/443/  table2.dat  (6Kb)

Sachkov M.E., Rastorguev A.S., Samus N.N., Gorynya N.A. Pis’ma Astron. Zh. 24, 443 (1998) The radii of 62 classical Cepheids.

J/PAZh/24/868/  table1.dat table5.dat  (16Kb)

Kiyaeva O.V., Tokovinin A.A., Kalinichenko O.A. Pis’ma Astron. Zh. 24, 868 (1998) The triple system ADS 9167.

J/PAZh/24/918/  table1.dat table5.dat  (16Kb)

Tokovinin A.A., Shatskii N.I., Magnitskii A.K.

Pis’ma Astron. Zh. 24, 918 (1998) ADS 9731: a new sextuple system.

J/PAZh/25/18/  object.dat table1.dat  (22Kb)

Sholukhova O.N., Fabrika S.N., Vlasyuk V.V. Pis’ma Astron. Zh. 25, 18 (1999) Spectroscopy of Hα-emission-line blue stars in M 33.

J/PAZh/25/30/  table1.dat  (11Kb)

Arkhipova V.P., Ikonnikova N.P., Noskova R.I., Sokol G.V., Esipov V.F., Klochkova V.G. Pis’ma Astron. Zh. 25, 30 (1999) IRAS 18062+2410: An early formation stage of a planetary nebula

J/PAZh/25/115/  table1.dat table3.dat  (23Kb)

Glushkova E.V., Batyrshinova V.M., Ibragimov M.A. Pis’ma Astron. Zh. 25, 115 (1999) A study of the open cluster NGC 6811.

J/PAZh/25/209/  table1.dat table2.dat  (13Kb)

Barannikov A.A. Pis’ma Astron. Zh. 25, 209 (1999) Long-term spectroscopic and photoelectric variability of the peculiar runaway O7I fp star HD 108

J/PAZh/25/276/  table4.dat  (34Kb)

Zakirov M.M. Pis’ma Astron. Zh. 25, 276 (1999) Close binary systems in star-forming regions: KV Cyg, V435 Cyg and V699 Cyg in OB association Cyg OB1

J/PAZh/25/403/  table.dat  (29Kb)

Lyuty V.M., Dorodhenko V.T. Pis’ma Astron. Zh., 25, 403 (1999) A new activity cycle of NGC 4151 : evidence for the existence of an accretion disk

J/PAZh/25/447/  table2.dat  (44Kb)

Gubanov A.G., Reshetnikov V.P. Pis’ma Astron. Zh. 25, 447 (1999) Identification of radio sources in clusters of galaxies. I.

J/PAZh/25/689/  table2.dat  (13Kb)

Rastorguev A.S., Glushkova E.V., Dambis A.K., Zabolotskikh M.V. Pis’ma Astron. Zh. 25, 689 (1999) Statistical parallaxes and kinematical parameters of classical Cepheids and young star clusters.

J/PAZh/25/764/  table1.dat  (17Kb)

Antokhina E.A., Kumsiashvili M.I. Pis’ma Astron. Zh. 25, 764 (1999) Parameters of the peculiar binary RY Sct as inferred from an analysis of new multicolor photometry.

J/PAZh/25/860/  table1.dat  (10Kb)

Taranova O.G., Shenavrin V.I. Pis’ma Astron. Zh. 25, 860 (1999) IR photometry of three carbon stars

J/PAZh/25/893/  table1.dat  (9Kb)

Belokon E.T., Babadzhanyants M.K. Pis’ma Astron. Zh. 25, 893 (1999) Spectral slope of the emission in ”fast” and ”slow” optical outbursts of the blazar 3C 345.

J/PAZh/26/13/  table.dat  (13Kb)

Lyuty V.M., Zaitseva G.V. Pis’ma Astron. Zh. 26, 13 (2000) One hundred years of observations of the Be star HDE 245770 (the X-ray binary A0535+26/V725 Tau): the end of an active phase?

J/PAZh/26/27/  table2.dat  (12Kb)

Karitskaya E.A., Goransky V.P., Grankin K.N., Mel’nikov S.Yu. Pis’ma Astron. Zh. 26, 27 (2000) Variability of Cyg X-1 (V1357) in 1995-1996 during the hard and soft spectral states.

J/PAZh/26/115/  table2.dat  (13Kb)

Klochkova V.G., Szczerba R., Panchuk V.E. Pis’ma Astron. Zh. 26, 115 (2000)

Optical spectrum of the infrared source IRAS 23304+6147.

J/PAZh/26/301/  refs.dat table1.dat  (17Kb)

Orlov V.V., Petrova A.V. Pis’ma Astron. Zh. 26, 301 (2000) Dynamical stability of triple stars.

J/PAZh/26/355/  table1.dat  (16Kb)

Emelyanov A.N., Aleksandrovich N.L., Sunyaev R.A. Pis’ma Astron. Zh. 26, 355 (2000) A Catalog of X-ray sources as observed by the TTM/COMIS telescope onboard the Mir-Kvant observatory in 1988-1998.

J/PAZh/26/465/  table2.dat table3.dat  (12Kb)

Klochkova V.G., Mishenina T.V., Panchuk V.E. Pis’ma Astron. Zh. 26, 465 (2000) Optical spectrum of the infrared source IRAS 20004+2955 (V1027 Cyg).

J/PAZh/26/510/  table2.dat table3.dat  (15Kb)

Klochkova V.G., Szczerba R., Panchuk V.E. Pis’ma Astron. Zh. 26, 510 (2000) Optical spectrum of the bipolar nebula AFGL 2688.

J/PAZh/26/525/  ubvri.dat  (26Kb)

Bondar N.I., Vitrichenko E.A., Zakirov M.M. Pis’ma Astron. Zh. 26, 525 (2000) A photometric study of the star V1016 Ori.

J/PAZh/26/535/  table1.dat  (8Kb)

Doroshenko V.T., Efimov Yu.S., Rosenbush A.E. Pis’ma Astron. Zh. 26, 535 (2000) UV Cas: Photometry, polarization, and spectrum near maximum light.

J/PAZh/26/695/  table.dat  (15Kb)

Taranova O.G., Shenavrin V.I. Pis’ma Astron. Zh. 26, 695 (2000) Infrared photometry of the symbiotic novae V1016 Cyg and HM Sge in 1978-1999.

J/PAZh/26/705/  table1.dat table2.dat table3.dat table4.dat table5.dat table6.dat  (20Kb)

Arkhipova V.P., Ikonnikova N.P., Noskova R.I., Sokol G.V. Pis’ma Astron. Zh. 26, 705 (2000) UBV observations of post-AGB supergiants and peculiarities of their photometric variability.

J/PAZh/27/83/  table.dat  (28Kb)

Doroshenko V.T., Lyuty V.M., Bochkarev N.G., Vlasyuk V.V., Mikhailov V.P., Spiridonova O.I., Grankin K.N., Ezhkova O.V., Mel’nikov S.Y. Pis’ma Astron. Zh. 27, 83 (2001) Nuclear activity of NGC 4151 as inferred from UBVRI observations in 1989-2000.

J/PAZh/27/132/  table3.dat  (65Kb)

Dambis A.K., Ratorguev A.S. Pis’ma Astron. Zh. 27, 132 (2001) Absolute Magnitudes and Kinematical Parameters of the Subsystem of RR Lyrae Type Variables

J/PAZh/27/170/  table1.dat table2.dat  (91Kb)

Kopylov A.I., Kopylova F.G. Pis’ma Astron. Zh. 27, 170 (2001) The Ursa Major supercluster of galaxies - I. The luminosity function.

J/PAZh/27/187/  table1.dat  (9Kb)

Arkhipova V.P., Ikonnikova N.P., Noskova R.I., Sokol G.V.,Shugarov S.Y. Pis’ma Astron. Zh. 27, 187 (2001) Light variations in the candidate for protoplanetary objects HD 179821 = V1427 Aql in 1899-1999.

J/PAZh/27/250/  table.dat  (9Kb)

Makarov D.I., Burenkov A.N., Tyurina N.V. Pis’ma Astron. Zh. 27, 250 (2001) Rotation curves for 135 edge-on galaxies.

J/PAZh/27/381/  table2.dat  (14Kb)

Vitrichenko E.A., Klochkova V.G.. Pis’ma Astron. Zh. 27, 381 (2001)

A spectroscopic study of the secondary star of BM Ori: Preliminary results.

J/PAZh/27/393/  table2.dat  (13Kb)

Taranova O.G., Shenavrin V.I. Pis’ma Astron. Zh. 27, 393 (2001) Infrared photometry of five long-period binaries.

J/PAZh/27/403/  table1.dat  (11Kb)

Kopylova F.G., Kopylov A.I. Pis’ma Astron. Zh. 27, 403 (2001) The Ursa Major supercluster of galaxies. II. Structure and peculiar velocities.

J/PAZh/27/442/  table1.dat  (23Kb)

Zakirov M.M. Pis’ma Astron. Zh. 27, 442 (2001) Close binary systems in star-forming regions: EQ Ori in the Ori I association.

J/PAZh/27/456/  table2.dat table3.dat table4.dat  (10Kb)

Kiyaeva O.V., Kiselev A.A., Polyakov E.V., Rafal’skii V.B. Pis’ma Astron. Zh. 27, 456 (2001) An astrometric study of the triple star ADS 48.

J/PAZh/27/632/  table2.dat  (21Kb)

Zakirov M.M. Pis’ma Astron. Zh. 27, 632 (2001) Close binary systems in star-forming regions: BZ Mon in the OB-association Mon OB2.

J/PAZh/27/841/  table1.dat table2.dat  (17Kb)

Arkhipova V.P., Ikonnikova N.P., Noskova R.I., Komissarova G.V., Klochkova V.G., Esipov V.F. Pis’ma Astron. Zh., 27, 841 (2001) Photometric variation and spectral peculiarities of protoplanetary nebulae LS II +34 26 = V1853 Cyg.

J/PAZh/27/942/  table2.dat  (17Kb)

Doroshenko V.T., Lyuty V.M., Bochkarev N.G., Merkulova N.I., Sergeev S.G., Vlasyuk V.V., Mikhailov V.P., Spiridonova O.I., Ezhkova O.V., Grankin K.N., Mel’nikov S.Yu. Pis’ma Astron. Zh. 27, 942 (2001) Nuclear activity of NGC 5548 from UBVRI observations during 1990-2000

J/PAZh/28/41/  table1.dat table2.dat table3.dat  (22Kb)

Masonkina L.I., Ryabtsev A.N., Ryabchikova T.A. Pis’ma Astron. Zh. 28, 41 (2002) Eu III oscillator strengths and Europium abundances in Ap stars.

J/PAZh/28/201/  refs.dat rem.dat vartype.txt vol1.dat vol1 pos.dat  (3.2 Mb)

Samus N.N., Goranskii V.P., Durlevich O.V., Zharova A.V., Kazarovets E.V., Pastukhova E.N., Hazen M.L., Tsvetkova T.M. Pis’ma Astron. Zh. 28, 201 (2002) An electronic version of Volume I of the General Catalogue of Variable Stars with improved coordinates.

J/PAZh/28/323/  table.dat  (10Kb)

Aleksandrovich N.L., Revnivtsev M.G., Aref’ev V.A., Sunyaev R.A., Skinner G.K. Pis’ma Astron. Zh. 28, 323 (2002) Long-Term-Mir-Kvant observations of the transient X-ray burster KS 1731-260.

J/PAZh/28/653/  objects.dat table1.dat table2.dat  (13Kb)

Sachkov M.E. Pis’ma Astron. Zh. 28, 653 (2002) Two Period-Radius Relations for Classical Cepheids: Determining the Pulsation Mode and the Distance Scale

J/PAZh/28/764/  table2.dat table3.dat  (10Kb)

Goranskii V.P., Kusakin A.V., Metlova N.V., Shugarov S.Yu., Barsukova E.A., Borisov N.V. Pis’ma Astron. Zh. 28, 764 (2002) Nova Monocetoris 2002 (V838 Mon) at early outburst stages

J/PAZh/29/253/  objects1.dat objects3.dat table1.dat table3.dat  (36Kb)

Bukhmastova Yu.L. Pis’ma Astron. Zh. 29, 253 (2003) Globular clusters as candidates for gravitational lenses to explain quasar-galaxy associations.

J/PAZh/29/571/  table.dat  (8Kb)

Malov I.F. Pis’ma Astron. Zh. 29, 571 (2004)

On the mechanism of X-ray emission from radio pulsars.

J/PAZh/29/858/  table1.dat table2.dat table3.dat table4.dat  (19Kb)

Arkhipova V.P., Ikonnikova N.P., Komissarova G.V., Noskova R.I., Shugarov S.Yu., Esipov V.F. Pis’ma Astron. Zh. 29, 858 (2003) FG Sagittae: new observations in the R CrB photometric phase

J/PAZh/30/37/  refs.dat table1.dat table2.dat table3.dat table4.dat table5.dat  (26Kb)

Tsvetkov D.Yu., Pavlyuk N.N. Pis’ma Astron. Zh. 30, 37 (2004) Observations of supernovae in the period 1997-1999.

J/PAZh/30/173/  table.dat  (11Kb)

Borkova T.V., Marsakov V.A. Pis’ma Astron. Zh. 30, 173 (2004) Stars of extragalactic origin in the solar neighborhood.

J/PAZh/30/243/  table1.dat table2.dat  (33Kb)

Hagen-Thorn V.A., Larionov V.M., Larionova E.G., Kudryavtseva N.A., Tikhonov A.V., Hagen-Thorn A.V., Arkharov A.A., Di Paola A., d’Alessio F. Pis’ma Astron. Zh. 30, 243 (2004) Optical and infrared monitoring of BL Lac in 1999-2001.

J/PAZh/30/297/  table1.dat table2.dat table3.dat table5.dat table6.dat table7.dat  (30Kb)

Arkhipova V.P., Esipov V.F., Ikonnikova N.P., Noskova R.I. Pis’ma Astron. Zh. 30, 297 (2004) Spectral and photometric evolution of the symbiotic Nova HM Sagittae in 1986-2002.

J/PAZh/30/430/  catalog.dat  (9Kb)

Revnivtsev M.G., Sunyaev R.A., Varshalovich D.A., Zheleznyakov V.V., Cherepashchuk A.M., Lutovinov A.A., Churazov E.M., Grebenev S.A., Gilfanov M.R. Pis’ma Astron. Zh. 30, 430 (2004) A hard X-ray survey of the galactic-center region with the IBIS telescope of the INTEGRAL observatory: a catalog of sources.

J/PAZh/30/589/  table.dat  (8Kb)

Molkov S.V., Cherepashchuk A.M., Lutovinov A.A., Revnivtsev M.G., Postnov K.A., Sunyaev R.A. Pis’ma Astron. Zh. 30, 589 (2004) A hard X-ray survey of the Sagittarius arm tangent with the IBIS telescope of the INTEGRAL observatory : a catalog of sources.

J/PAZh/30/605/  table1.dat table2.dat  (16Kb)

Taranova O.G., Shenavrin V.I. Pis’ma Astron. Zh. 30, 605 (2004) JHKLM photometry for carbon stars.

J/PAZh/30/692/  notes.txt table.dat  (14Kb)

Denissenko D.V. Pis’ma Astron. Zh. 30, 692 (2004) Occultations of HIP and UCAC2 stars downto 15m by large TNO 2004-2014.

J/PAZh/30/855/  table1.dat  (7Kb)

Arkhipova V.P., Ikonnikova N.P., Noskova R.I., Komissarova G.V., Esipov V.F. Pis’ma Astron. Zh. 30, 855 (2004) Photometric variability and spectrum of the post-AGB candidate IRAS 19200+3457.

J/PAZh/30/946/  table.dat  (24Kb)

Simonia I.A. Pis’ma Astron. Zh. 30, 946 (2004) Unidentified cometary emission lines as the photoluminescence of frozen hydrocarbon particles.

J/PAZh/31/116/  table.dat  (8Kb)

Zaitseva G.V. Pis’ma Astron. Zh. 31, 116 (2005) Further photometric observations of the X-ray binary A0535+26/V725 Tau: a new activity cycle.

J/PAZh/31/332/  table1.dat  (46Kb)

Karachentsev I.D., Kutkin A.M. Pis’ma Astron. Zh. 31, 332 (2005) Infrared Luminosities of Local-Volume Galaxies.

J/PAZh/31/437/  objects.dat refs.dat table.dat  (19Kb)

Ryabchikova T.A. Pis’ma Astron. Zh. 31, 437 (2005) Temperature behavior of elemental abundances in the atmospheres of magnetic peculiar stars.

J/PAZh/31/563/  table1.dat table2.dat table3.dat  (20Kb)

Mitronova S.N., Huchtmeier W.K., Karachentsev I.D., Karachentseva V.E., Kudrya Yu.N. Pis’ma Astron. Zh. 31, 563 (2005) HI observations of flat galaxies.

J/PAZh/32/48/  table1.dat  (10Kb)

Arkhipova V.P., Ikonnikova N.P., Komissarova G.V., Esipov V.F. Pis’ma Astron. Zh. 32, 48 (2006) Study of the IR-Excess Supergiant HD 331319.

J/PAZh/32/508/  table2.dat  (13Kb)

Chelovekov I.V., Grebenev S.A., Sunyaev R.A. Pis’ma Astron. Zh. 32, 508 (2006) Hard X-ray bursts detected by the IBIS telescope onboard the INTEGRAL Observatory in 2003-2004.

J/PAZh/32/628/  table1.dat table2.dat  (228Kb)

Arkharov A.A., Ruban E.V., Novikov V.V., Yagudin L.I., Puzakova T.Yu., Zalles R. Pis’ma Astron. Zh. 32, 628 (2006) Spectroscopic observations of southern stars in the wavelength range 320-730nm.

J/PAZh/32/662/  table1.dat  (9Kb)

Arkhipova V.P., Ikonnikova N.P., Komissarova G.V., Esipov V.F. Pis’ma Astron. Zh. 32, 662 (2006) Spectroscopic and Photometric Variability of HD 51585 As a Manifestation of Stellar Wind.

J/PAZh/32/672/ table1.dat table2.dat table3.dat table4.dat table5a.dat table5b.dat table5c.dat (1.0 Mb) Ruban E.V., Alekseeva A.A., Arkharov A.A., Hagen-Thorn E.I., Galkin V.D., Nikanorova I.N., Novikov V.V., Pakhomov V.P., Puzakova T.Yu. Pis’ma Astron. Zh. 32, 672 (2006) Spectrophotometric observations of variable stars.

J/PAZh/32/737/  table3.dat table7.dat  (11Kb)

Arkhipova V.P., Klochkova V.G., Chentsov E.L., Esipov V.F., Ikonnikova N.P., Komissarova G.V. Pis’ma Astron. Zh. 32, 737 (2006) Spectroscopy and photometry of the protoplanetary nebula candidate StHa62 = IRAS 07171+1823.

J/PAZh/32/866/  table1.dat table2.dat  (12Kb)

Taranova O.G., Shenavrin V.I. Pis’ma Astron. Zh. 32, 866 (2006) JHKLM photometry for 22 hot stars.

J/PAZh/32/906/  table1.dat table2.dat  (9Kb)

Bobylev V.V. Pis’ma Astron. Zh. 32, 906 (2006) Kinematics of Gould Belt Based on Open Clusters.

J/PAZh/33/143/  table1.dat table2.dat table3.dat table6.dat  (31Kb)

Lapinov A.V., Golubiatnikov G.Yu., Markov V.N., Guarnieri A. Pis’ma Astron. Zh. 33, 143 (2007) Laboratory studies of the HNCO molecular spectrum for precise spectroscopy of dark clouds.

J/PAZh/33/403/  table1.dat table2.dat table4.dat  (25Kb)

Bukhmastova Yu.L. Pis’ma Astron. Zh. 33, 403 (2007) Quasars Lensed by Globular Clusters of Spiral and Elliptical Galaxies.

J/PAZh/33/679/  table1.dat table4.dat  (18Kb)

Arkhipova V.P., Esipov V.F., Ikonnikova N.P., Komissarova G.V., Noskova R.I. Pis’ma Astron. Zh. 33, 604 (2007) Variability and rapid evolution of the protoplanetary object IRAS 18062+2410 = V886 Her.

J/other/PNAOJ/2.561/ table1 table2 (34Kb) Miura N., Baba N., Ni-Ino M., Ohtsubo J., Noguchi M., Isobe S. Publ. Nat. Astron. Obs. Japan, 2, 561 (1992) Speckle observations of visual and spectroscopic binaries. IV.

J/other/ChJAA/3.49/ refs.dat table1.dat (66Kb) Xu Y., Zheng X.-W., Jiang D.-R. Chinese J. Astron. Astrophys., 3, 49-68 (2003) Statistical properties of 6.7GHz methanol maser sources.

J/other/NewA/5.137/ table1.dat (11Kb) Greiner J. New Astronomy, 5, 137 (2000) Catalog of supersoft X-ray sources

J/other/PNAOJ/5.75/ a0097.dat a0477.dat a0481.dat a0535.dat a0754.dat a0814.dat a0815.dat a0825.dat a0886.dat a0966.dat notes.dat (63Kb) Miyauchi-Isobe N., Maehara H. Publ. Natl. Astron. Obs. Japan, 5, 75-97 (1998) The Second Kiso Survey for Ultraviolet-Excess Galaxies. I (KUG2-1)

J/other/PNAOJ/6.1/ a0060.dat a0134.dat a0485.dat a0497.dat a0504.dat a0562.dat a0634.dat a0707.dat a0786.dat a0860.dat notes.dat (103Kb) Miyauchi-Isobe N., Maehara H. Publ. Natl. Astron. Obs. Japan, 6, 1-39 (2000) The Second Kiso Survey for Ultraviolet-Excess Galaxies. II (KUG2-2)

J/other/PNAOJ/6.107/ clusters notes.dat tables.dat (68Kb) Miyauchi-Isobe N., Maehara H. Publ. Nat. Astron. Obs. Japan, 6, 107 (2002) The Second Kiso Survey for ultraviolet-excess galaxies. III.

J/other/ChJAA/6.237/ refs.dat table1.dat (104Kb) You X.-P., Han J.-L. Chinese J. Astron. Astrophys., 6, 237-246 (2006) Circular polarization in pulsar integrated profiles: updates.

J/other/ChJAA/6.435/ table5.dat (28Kb) Zhao J.-L., Chen L., Wen W. Chinese J. Astron. Astrophys., 6, 435 (2006) Membership and segregation effects in the young open cluster NGC 6530.

J/other/PASAu/6.471/ table1 table2 (27Kb) McGee R.X., Newton L.M. Proc. Astron. Soc. Australia, 6, 471-500 (1986) HI profiles in the bridge region of the Magellanic Clouds

J/other/PNAOJ/7.37/ notes.dat table4.dat (25Kb) Miyauchi-Isobe N., Maehara H. Publ. Nat. Astron. Obs. Japan, 7, 37-52 (2003) The Second Kiso Survey for ultraviolet-excess galaxies. IV.

J/other/NewA/8.645/ table1.dat (11Kb) Sharaf M.A. New Astron., 8, 645-653 (2003) Relation between the apparent magnitude and the parallax for Hyades stars.

J/other/NewA/9.509/ table1.dat (24Kb) Zhang P., Chen P.S., He J.H. New Astronomy, 9, 509-524 (2004) A new approach to IRAS observations of Be stars.

J/other/PASAu/10.140/catalog.dat (100Kb) White G.L. Proc. Astron. Soc. Australia, 10, 140 (1992) A catalogue of ’Southern’ radio sources with accurate positions

J/other/PASAu/10.310/table1.dat (32Kb) Duncan R.A., White G.L., Wark R., Reynolds J.E., Jauncey D.L., Norris R.P., Taaffe L., Savage A. Proc. Astron. Soc. Australia, 10, 310 (1993) A survey of 526 southern flat-spectrum radio sources with the Parkes-Tidbinbilla interferometer

J/other/NewA/10.325/ table1.dat (158Kb) Zhang P., Chen P.S., Yang H.T. New Astronomy, 10, 325 (2005) 2MASS observations of Be stars.

J/other/RoAJ/10.51/ table1.dat (13Kb) Bocsa G. Romanian Astronomical Journal, 10, 51 (2000) Contributions of the Bucharest observatory to the CONFOR program.

J/other/NewA/11.262/ table1.dat (1.0 Mb) Piatti A.E., Claria J.J., Ahumada A.V. New Astronomy, 11, 262-272 (2006) New fundamental parameters for the inner disk open cluster Lynga 11.

J/other/NewA/11.27/ table1.dat (80Kb) Chen P.S., Fu H.W., Gao Y.F. New Astronomy, 11, 27-42 (2005) 2MASS observation of BL Lac objects.

J/other/NewA/11.293/ table1.dat (12Kb) Ann H.B., Park J.-C. New Astronomy, 11, 293-305 (2006) Warped disks in spiral galaxies.

J/other/NewA/11.557/ table1.dat table2.dat table3.dat (25Kb) Chen P.S., Shan H.G., Gao Y.F. New Astronomy, 11, 557-566 (2006) On the difference between the X-ray selected and the radio selected BL Lac objects in the near infrared.

J/other/KFNT/12.20/ catalog.dat (24Kb) Zakhozhaj V.A., Shaparenko E.F. Kinematika Fiz. Nebesn. Tel., 12, part no 2, 20 (1996) The catalogue of nearby stars metallicities.

J/other/NewA/12.446/ refs.dat table3.dat (14Kb) Ak T., Bilir S., Ak S., Retter A. New Astronomy, 12, 446-453 (2007) A new absolute magnitude calibration with 2MASS for cataclysmic variables.

J/other/NewA/12.461/ table2.dat (93Kb) Giorgi E., Vazquez R.A., Solivella G.R., Orellana R.B., Nunez J. New Astron., 12, 461-470 (2007) Revealing the parameters of the open cluster Ruprecht 58.

J/other/SAAOC/14.11/table1 (27Kb) Kilkenny D., Laing J.D. South Afr. Astron. Obs. Circ. 14, 11 (1990) UBVRI observations of stars in the Guide Star Photometric Catalogue

J/other/AuJPA/14.77/ table1.dat (26Kb) Shaver P.A., Goss W.M. Australian J. Phys. Astrophys. Suppl., 14, 77 (1970) Galactic radio sources. II. High resolution 408 MHz observations

J/other/SAAOC/15.30/table1 table1a table2 table2b table3 (21Kb) Cousins A.W.J. South Afr. Astron. Obs. Circ. 15, 30 (1993) DDO photometry of E-region stars and equatorial standards (Paper I).

J/other/SAAOC/15.40/table1 table2 table3 (48Kb) Kilkenny D., Whittet D.C.B. South Afr. Astron. Obs. Circ. 15, 40 (1993) A Study of Reddened Luminous Stars in the Southern Milky Way I. uvby Photometry

J/other/KFNT/15.483/ jovsat.dat (7Kb) Ledovskaya I.V., Jockers K., Karpov N.V., Sergeev A.V. Kinematika Fiz. Nebesn. Tel., 15, 483 (1999) Astrometric CCD observations of Jupiter’s inner satellites Thebe and Amalthea and Saturn’s satellite Phoebe in 1998 opposition.

J/other/SAAOC/15.53/table1 table2 (27Kb) Kilkenny D. South Afr. Astron. Obs. Circ. 15, 53 (1993) A Study of Reddened Luminous Stars in the Southern Milky Way II. Hbeta Photometry, Reddening and Distances J/other/SAAOC/15.65/table1 table2 (17Kb)

Kilkenny D., van Wyk F., Marang F., Spencer Jones J.H., Whittet D.C.B. South Afr. Astron. Obs. Circ. 15, 65 (1993) A Study of Reddened Luminous Stars in the Southem Milky Way III. VRI Photometry

J/other/SAAOC/15.73/table1 (14Kb) Spencer Jones J.H., Whittet D.C.B., He L. South Afr. Astron. Obs. Circ. 15, 73 (1993) A Study of Reddened Luminolls Stars in the Southern Milky Way IV. JHK Photometry

J/other/SAAOC/15.78/table1 (29Kb) Kilkenny D., Barrett P.E., Marang F., van Wyk F., Roberts G., Winkler H. South Afr. Astron. Obs. Circ. 15, 78 (1993) UBV Photometry of HD Stars in the Fields of Selected Cataclysmic Variables

J/other/PASA/16.113/ table2.dat table3.dat (24Kb) Drinkwater M.J., Proust D., Parker Q.A., Quintana H., Slezak E. Publications - Astron. Soc. Australia, 16, 113 (1999) New structure in the Shapley supercluster

J/other/PBeiO/16.18/ table (42Kb) Luo D., Zhang B. Publ. Beijing Astron. Obs. 16, 18 (1990) Physical data of the fundamental stars

J/other/PBeiO/17.105/ table1 (36Kb) Lu L.-Z. Publ. Beijing Astron. Obs., 17, 105 (1990) Third preliminary catalogue of stars’ right ascension observed with the photoelectric transit instrument (PPCP3)

J/other/PBeiO/17.120/ table1 (26Kb) Wang R., Lu L. Publ. Beijing Astron. Obs., 17, 120 (1991) Fourth preliminary catalogue of stars right ascensions observed with photoelectric transit instrument (PPCP4)

J/other/PASA/17.92/ mnp01.ps mnp02.ps mnp03.ps mnp04.ps mnp05.ps table1.dat (402Kb) Otrupcek R.E., Hartley M., Wang J.-S. Publications - Astron. Soc. Australia, 17, 92 (2000) Catalogue of J = 1-0 CO emission towards southern dark clouds.

J/other/PASA/18.232/ table2.dat table3.dat (14Kb) Mamon G.A., Parker Q.A., Proust D. Publications - Astron. Soc. Australia, 18, 232 (2001) FLAIR II spectroscopy of two DENIS J band galaxy samples.

J/other/PBeiO/18.7/ table1 table2 table3 table4 (17Kb) Liu Y.-L. Publ. Beijing Astron. Obs., 18, 7 (1991) Far infrared Seyfert galaxies and their confusing point sources

J/other/PASA/19.455/ notes.dat refflux.dat refnhi.dat table1.dat table2.dat (95Kb) Curran S.J., Webb J.K., Murphy M.T., Bandiera R., Corbelli E., Flambaum V.V. Publications - Astron. Soc. Australia, 19, 455 (2002) A catalogue of damped Lyman alpha absorption systems and radio flux densities of the background quasars.

J/other/AcApS/20.366/table1.dat table2.dat (23Kb) Chen Y., Han J.-L., He X.-T. Acta Astrophys. Sinica, 20, 366 (2000) The radio quasars identified from the NVSS catalogue.

J/other/KFNT/23.102/ redclump.dat (7.1 Mb) Rybka S.P. Kinematika Fiz. Nebesn. Tel., 23, 102 (2007) A catalogue of candidate Red Clump stars in the Tycho-2.

J/other/PZ/23.141/  table.dat  (24Kb)

Kazarovets E.V. Perem. Zvezdy, 23, 141 (1993) Coordinates and cross-identification of flare stars in Pleiades

J/other/PASA/23.154/ aqpsc.dat eecet.dat xyleo.dat (47Kb)

Djurasevic G., Dimitrov D., Arbutina B., Albayrak B., Selam S.O., Atanackovic-Vukmanovic O. Publ. Astron. Soc. Australia, 23, 154 (2006) A photometric study of the contact binaries: XY Leo, EE Cet and AQ Psc.

J/other/PBeiO/24.71/ catalog.dat (117Kb) Lu L., Manrique W.T., Perdomo R. Publ. Beijing Astron. Obs., 24, 71 (1994) The position corrections of 1400 stars observed with PA II in San Juan

J/other/JApA/25.143/ appena.dat appenc.dat (57Kb) Mohan R., Dwarakanath K.S., Srinivasan G. J. Astrophys. Astron., 25, 143-183 (2004) A high galactic latitude HI 21 cm-line absorption survey using the GMRT. I. Observations and spectra.

J/other/APh/26.282/ tableb1.dat tableb2.dat tableb3.dat tableb4.dat tableb5.dat tableb6.dat tableb7.dat tableb8.dat tableb9.dat tableb10.dat tableb11.dat (33Kb) Achterberg A., Ackermann M., Adams J., Ahrens J., Atlee D.W., Bahcall J.N., Bair X., Baret B., Bartelt M., Barwick S.W., Bay R., Beattie K., Becka T., Becker J.K., Becker K.-H., Berghaus P., Berley D., Bernardini E., Bertrand D., Besson D.Z., Blaufuss E., Boersma D.J., Bohm C., Boeser S., Botner O., Bouchta A., Braun J., Burgess C., Burgess T., Castermans T., Chirkin D., Clem J., Collin B., Conrad J., Cooley J., Cowen D.F., D’agostino M.V., Davour A., Day C.T., De Clercq C., Desiati P., Deyoung T., Dreyer J., Duvoort M.R., Edwards W.R., Ehrlich R., Ellsworth R.W., Evenson P.A., Fazely A.R., Feser T., Filimonov K., Gaisser T.K., Gallagher J., Ganugapati R., Geenen H., Gerhardt L., Goldschmidt A., Goodman J.A., Greene M.G., Grullon S., Gross A., Gunasingha R.M., Hallgren A., Halzen F., Han K., Hanson K., Hardtke D., Hardtke R., Harenberg T., Hart J.E., Hauschildt T., Hays D., Heise J., Helbing K., Hellwig M., Herquet P., Hill G.C., Hodges J., Hoffman K.D., Hoshina K., Hubert D., Hughey B., Hulth P.O., Hultqvist K., Hundertmark S., Ishihara A., Jacobsen J., Japaridze G.S., Jones A., Joseph J.M., Kampert K.-H., Karle A., Kawai H., Kelley J.L., Kestel M., Kitamura N., Klein S.R., Klepser S., Kohnen G., Kolanoski H., Koepke L., Krasberg M., Kuehn K., Landsman H., Lang R., Leich H., Leuthold M., Liubarsky I., Lundbert J., Madsen J., Mase K., Matis H.S., McCauley T., McParland C.P., Meli A., Messarius T., Meszaros P., Minor R.H., Miocinovic P., Miyamoto H., Mokhtarani A., Montaruli T., Morey A., Morse R., Movit S.M., Munich K., Nahnhauer R., Nam J.W., Niessen P., Nygren D.R., Oegelman H., Olbrechts PH., Olivas A., Patton S., Pena-Garay C., Perez de los Heros C., Pieloth D., Pohl A.C., Porrata R., Pretz J., Price P.B., Przybylski G.T., Rawlins K., Razzaque S., Refflinghaus F., Resconi E., Rhode W., Ribordy M., Richter S., Rizzo A., Robbins S., Rott C., Rutledge D., Sander H.-G., Schlenstedt S., Schneider D., Seckel D., Seo S.H., Seunarine S., Silvestri A., Smith A.J., Solarz M., Song C., Sopher J.E., Spiczak G.M., Spiering C., Stamatikos M., Stanev T., Steffen P., Steele D., Stezelberger T., Stokstad R.G., Stoufer M.C., Stoyanov S., Sulanke K.-H., Sullivan G.W., Sumner T.J., Taboada I., Tarasova O., Tepe A., Thollander L., Tilav S., Toale P.A., Turcan D., Van Eijndhoven T.J., Vandenbroucke J., Voigt B., Wagner W., Walck C., Waldmann H., Walter M., Wang Y.-R., Wendt C., Wiesbusch C.H., Wikstroem G., Williams D.R., Wischnewski R., Wissing H., Woschnagg K., Xu X.W., Yodh G., Yoshida S., Zornoza J.D., Biermann P.L. Astroparticle Phys., 26, 282-300 (2006) On the selection of AGN neutrino source candidates for a source stacking analysis with neutrino telescopes.

J/other/APh/26.367/ appen.dat (8Kb) Besson D., Razzaque S., Adams J., Harris P. Astroparticle Phys., 26, 367-377 (2007) Limits on the transient ultra-high energy neutrino flux from gamma-ray bursts (GRB) derived from RICE data.

J/other/ChA+A/28.308/ table1.dat table2.dat table3.dat (11Kb) Fang M., Yao Y.-Q. Chinese Astron., 28, 308-322 (2004) Near infrared observations of the star formation region GGD 12-15.

J/other/BASI/29.519/ table.dat (226Kb) Sagar R., Naidu B.N., Mohan V. Bull. Astron. Soc. India, 29, 519 (2001) A deep VI CCD photometric study of the neglected galactic star cluster NGC 6631.

J/other/ApL/30.1/  catalog.dat table2.dat  (78Kb)

Reduzzi L., Rampazzo R. Astrophys. Lett., 30, 1 (1995) Candidates for a southern extension of the Karachentsev catalogue of isolated pairs of galaxies

J/other/BASI/30.911/ table6.dat table7.dat table8.dat table9.dat tables.ps tables.tex (286Kb) Sharma M., Prabhu T.P. Bull. Astron. Soc. India, 30, 911 (2002)

Photometric catalogs of four EMSS poor clusters of galaxies.

J/other/Ap/31.714/  table1.dat  (11Kb)

Kurtanidze O.M., Nikolashvili M.G. Astrophysics, 31, 714 (1989) A low dispersion sky spectral survey for revealing faint carbon stars. IV. Region 90^{∘}=l =115^{∘}, -5^{∘}=b =+5^{∘}

J/other/RMxAA/32.143/ table2 (16Kb) Ruelas-Mayorga A., Garcia-Ruiz G. Rev. Mex. Astron. Astrofis., 32, 143 (1996) Stromgren photometry of stars in Baade’s Window. I. The data.

J/other/BASI/32.295/ table3.dat table4.dat table5.dat (28Kb) Sujatha S., Babu G.S.D., Ananthamurthy S. Bull. Astron. Soc. India, 32, 295-309 (2004) UBVRI CCD photometric studies of open clusters Berkeley 15, Czernik 18 and NGC 2401.

J/other/AISAO/32.31/ tab3-5 tab6-8 (44Kb) Afanasiev V.L., Vlasyuk V.V., Dodonov J.N., Lorens H., Terebbizh V.Y. Astrofizik. Issledovanija, 32, 31 (1990) Spectral and photometrical investigation of objects in selected fields with the 6-m telescope

J/other/PBeiO/32.87/ table1.dat table2.dat (15Kb) Manrique W.T., Lu L.Z., Podesta R.C., Wang Z.Z. Publ. Beijing Astron. Obs., 32, 87 (1998) San Juan radio star catalogue and comparison with Hipparcos catalogue

J/other/CoSka/33.159/ table1.dat table2.dat (38Kb) Neslusan L. Contr. Astron. Obs. Skalnate Pleso, 33, 159-172 (2003) Astrometry of minor planets made at the Skalnate Pleso Observatory in the year 2002.

J/other/RMxAA/33.83/table1.dat table3.dat (35Kb) Stock J., Stock J. Rev. Mex. Astron. Astrofis., 33, 83 (1997) Positions and proper motions in the area of the open cluster NGC 129

J/other/BASI/35.53/ table23.dat (7Kb) Sharma S., Ogura K., Pandey A.K., Ojha D.K., Bhatt B.C. Bull. Astron. Soc. India, 35, 53-63 (2007) Near-infrared JH photometry of young open cluster Be 59.

J/other/RMxAA/37.115/ notes.dat table1.dat table2.dat table3.dat (120Kb) Kimeswenger S. Rev. Mex. Astron. Astrofis., 37, 115-145 (2001) Optical coordinates of southern planetary nebulae.

J/other/RMxAA/38.129/ table3a.dat table3b.dat table3c.dat table3d.dat table4a.dat table4b.dat table4c.dat table4d.dat table6a.dat table6b.dat table6c.dat table6d.dat table7a.dat table7b.dat table7c.dat table7d.dat (38Kb) Stock J., Stock M., Garcia J., Sanchez N. Rev. Mex. Astron. Astrofis., 38, 129 (2002) Quantitative Stellar Spectral Classification. II. Early Stars.

J/other/RMxAA/38.141/ table1.dat (49Kb) Moreno-Corral M.A., Chavarria-K C., de Lara E. Rev. Mex. Astron. Astrofis., 38, 141 (2002) CCD imagery, uvby-beta photometry, and the physical parameters of Haffner 19.

J/other/BSAO/38.152/ table1.dat table2.dat (57Kb) Glagolevskij Yu.V. Bull. Special Astrophys. Obs., 38, 152 (1994) A new list of effective temperatures of chemically peculiar stars. II.

J/other/RMxAA/39.39/table4.dat table5.dat table6.dat (162Kb) Ahumada J.A. Rev. Mex. Astron. Astrofis. 39, 39 (2003) CCD Photometry of the Open Clusters NGC 2658, NGC 2849, and NGC 3247

J/other/RMxAA/39.89/stars.dat table3.dat table4.dat (29Kb) Vazquez R.A., Giorgi E.E., Brusasco M.A., Baume G., Solivella G.R. Rev. Mex. Astron. Astrofis., 39, 89-105 (2003)

Lynga 1, a small open cluster containing a red-supergiant member.

J/other/BSAO/40.128/ table3.dat (48Kb) Bursov N.N Bull. Special Astrophys. Obs., 40, 128 (1996) Spectra of radio sources of the program ”Cold” at the RATAN-600 radiotelescope

J/other/BSAO/40.5/ notes.dat table1.dat table2.dat table3.dat (96Kb) Parijskij Yu.N., Goss W.M., Kopylov A.I., Soboleva N.S., Temirova A.V., Verkhodanov O.V., Zhelenkova O.P., Naugolnaya M.N. Bull. Special Astrophys. Obs., 40, 5 (1996) Investigation of RATAN-600 RC radio sources

J/other/Ap/41.356/  notes-a.dat notes.dat table1-a.dat table1.dat  (41Kb)

Gigoyan K.S., Hambaryan V.V., Azzopardi M. Astrophys., 41, 356-366 (1998) First Byurakan spectral sky survey. Stars of late spectral types. -11^{∘}δ-7^{∘}belt.

J/other/RMxAA/42.75/table1.dat table2.dat table4.dat (68Kb) Arellano Ferro A., Garcia Lugo G., Rosenzweig P. Rev. Mex. Astron. Astrofis., 42, 75-87 (2006) CCD Photometry of the Globular Cluster M15: RR Lyrae Fourier Decomposition and Physical Parameters.

J/other/ARA+A/43.195/ refs.dat table3.dat (19Kb) Kirkpatrick J.D. Annual Rev. Astron. Astrophys., 43, 195 (2005) New spectral types L and T.

J/other/Ap/44.328/  notes.dat notes a.dat table1.dat table1 a.dat  (22Kb)

Gigoyan K.S., Abrahamyan H.V., Azzopardi M., Russeil D. Astrophysics, 44, 328 (2001) First Byurakan spectral sky survey. Stars of late spectral types. XII. Zone -3^{∘}δ+1^{∘}.

J/other/Ap/44.335/  notes.dat table23.dat  (32Kb)

Kazarian M.A., Petrosian G.V. Astrophysics, 44, 335 (2001) Classification of red stars with spectral absorption bands observed in the near-infrared range.

J/other/Ap/44.88/  notes.dat notes a.dat table1.dat table1 a.dat  (27Kb)

Gigoyan K.S., Abrahamyan H.V., Azzopardi M., Russeil D. Astrophysics, 44, 88 (2001) First Byurakan Spectral Sky Survey. Stars of late spectral types. XI. Zone -7^{∘}δ-3^{∘}.

J/other/BSAO/45.136/ table1.dat (26Kb) Ugolkova L.S. Bull. Special Astrophys. Obs., 45, 136 (1998) Optical identification of Zelenchuk survey radio sources from 9h to 12h in right ascension and between 0^{∘}and 8^{∘}in declination. II. Some optical characteristics of identified radio sources.

J/other/Ap/45.322/  notes.dat table1.dat  (20Kb)

Gigoyan K.S., Abrahamyan H.V., Azzopardi M., Russeil D. Astrophysics, 45, 322 (2002) First Byurakan Spectral Sky Survey types. XIII. Zone +1^{∘}=δ=13^{∘}.

J/other/Ap/46.234/  refs.dat table1.dat table2.dat table3.dat table4.dat table5.dat  (66Kb)

Kudryavtsev D.O., Romanyuk I.I. Astrophysics, 46, 234 (2003) Spatial distribution of magnetic CP stars

J/other/BSAO/46.28/ table2.dat table3.dat (47Kb) Mingaliev M.G., Botashev A.M., Stolyarov V.A. Bull. Special Astrophys. Obs., 46, 28 (1998) Simultaneous spectra of complete sample of sources from the PMN survey

J/other/Ap/46.46/  notes.dat refs.dat table1.dat  (17Kb)

Abrahamyan G.V., Sinamyan P.K., Gigoyan K.S. Astrophysics Ap, 46.46 (2003) First Byurakan Spectral Sky Survey. Blue stellar objects. XII. The region -3^{∘}δ+1^{∘}.

J/other/Ap/46.475/  notes.dat table1.dat table2.dat  (20Kb)

Gigoyan K.S., Abrahamyan H.V., Azzopardi M., Mauron N., Russeil D., Sinamyan P. Astrophysics, 46, 475-488 (2003)

First Byurakan Spectral Sky Survey. Stars of late-spectral types. XIV. Zone +13 degrees = delta = +33 degrees.

J/other/Ap/47.305/  table1.dat  (14Kb)

Eritsian M.A. Astrophysics, 47, 305-312 (2004) Circular polarization of starlight.

J/other/Ap/48.304/  fields.dat refs.dat table2.dat table3.dat  (30Kb)

Doroshenko V.T., Sergeev S.G., Merkulova N.I., Sergeeva E.A., Golubinsky Y.V., Pronik V.I., Okhmat N.N. Astrophysics, 48, 304-320 (2005) BVRI CCD-Photometry of comparison stars in the neighborhood of galaxies with active nuclei. II.

J/other/Ap/48.344/  table1.dat  (13Kb)

Kazarian M.A., Petrosian G.V. Astrophysics, 48, 344-354 (2005) New galaxies with UV excess. VI.

J/other/BSAO/50.5/ table2.dat (105Kb) Karachentsev I.D., Karachentseva V.E., Kudrya Y.N., Makarov D.I., Parnovsky S.L. Bull. Special Astrophys. Obs., 50, 5 (2000) A list of peculiar velocities of RFGC galaxies.

J/other/P+SS/53.925/ cl/* lc/* objects.dat (97Kb) Szekely P., Kiss L.L., Szabo Gy.M., Sarneczky K., Csak B., Varadi M., Meszaros Sz. Planetary and Space Science, 53, 925-936 (2005) CCD photometry of 23 minor planets.

J/other/BSAO/57.5/ catalog.dat (1.5 Mb) Mitronova S.N., Karachentsev I.D., Karachentseva V.E., Jarrett T.H., Kudrya Yu.N. Bulletin of the Special Astrophysical Observatory, 57, 5-163 (2004) The 2MASS-selected Flat Galaxy Catalog

J/other/Obs/116.382/ obs.dat (29Kb) Miller R., Osborn W. The Observatory 116, page (1996) BVRI Photometry of Spectroscopic Binaries

J/other/Sci/292.2290/ table1.dat (15Kb) Grindlay J.E., Heinke C., Edmonds P.D., Murray S.S. Science, 292, 2290 (2001) High-resolution X-ray imaging of a globular cluster core: compact binaries in 47 Tuc

J/other/Sci/292.698/ ohdhsall.dat ohdhsobs.dat table1.dat (570Kb) Oppenheimer B.R., Hambly N.C., Digby A.P., Hodgkin S.T., Saumon D. Science 292, 698 (2001) Direct detection of Galactic halo dark matter

J/other/GCN/1070.1/ hete.dat hete2001.dat hete2002.dat (150Kb) Ricker G., Lamb D., Woosley S., Vanderspek R., Crew G., Doty J., Monnelly G., Villasenor J., Butler N., Cline T., Jernigan J.G., Levine A., Martel F., Morgan E., Pizzichini G., Prigozhin G., Kawai N., Matsuoka M., Shirasaki Y., Tamagawa T., Torii K., Sakamoto T., Yoshida A., Fenimore E., Galassi M., Graziani C., Atteia J.-L., Boer M., Olive J.-F., Dezalay J.-P., Hurley K. GRB Circular Network, 1070, 1 (2001) SGR1806-20: A Second, Intense Burst Localized by HETE.

J/other/IBVS/4431/ notes.dat table1.dat (13Kb) Skiff B. IAU Inform. Bull. Var. Stars, 4431, 1 (1997) Identification of the Margoni-Stagni variables

J/other/IBVS/4433/ notes.dat table1.dat (15Kb) Skiff B.A. IAU Inform. Bull. Var. Stars, 4433, 1 (1997) Identification of the Rosino-Guzzi variables in Sagitta

J/other/IBVS/4441/ notes.dat table1.dat table2.dat (11Kb) Skiff B.A. IAU Inform. Bull. Var. Stars, 4441, 1 (1997) Identification of variables near NGC 7635.

J/other/IBVS/4459/ notes.dat table1.dat table2.dat (14Kb)

Skiff B.A. IAU Inform. Bull. Var. Stars, 4459, 1 (1997) Identifications for Baade’s variables in Sagitta and Cygnus

J/other/IBVS/4678/ notes.dat table1.dat (15Kb) Skiff B.A. IAU Inform. Bull. Var. Stars, 4678, 1 (1999) Identifications for the Haro-Chavira ”infrared stars”.

J/other/IBVS/5185/ notes.dat table1.dat (14Kb) Skiff B.A. IAU Inform. Bull. Var. Stars, 5185, 1 (2001) Coordinates and identifications for Dolidze S, C, and MS stars

J/other/IBVS/5187/ notes.dat table1.dat (13Kb) Skiff B.A IAU Inform. Bull. Var. Stars, 5187 (2001) Coordinates and identifications for Rosino’s red variables near NGC 6749.

J/other/IBVS/5296/ crossid.dat notes.dat table1.dat table2.dat (72Kb) Agerer F., Hubscher J. IAU Inform. Bull. Var. Stars, 5296, 1 (2002) Photoelectric minima of selected eclipsing binaries and maxima of pulsating stars.

J/other/IBVS/5298/ notes.dat ori/* refs1.dat refs2.dat refs3.dat refs4.dat refs5.dat table1.dat table2.dat table3.dat table4.dat table5.dat (284Kb) Webbink R.F., Hazen M.L., Hoffleit D. IAU Inform. Bull. Var. Stars, 5298, 1 (2002) Coordinates and identifications of Harvard variables

J/other/IBVS/5438/ stars.dat table1.dat (55Kb) Diethelm R. IAU Inform. Bull. Var. Stars, 5438, 1 (2003) 162-nd list of minima timings of eclipsing binaries by BBSAG observers.

J/other/IBVS/5480/ notes.dat refs.dat table1.dat (19Kb) Otero S.A. IAU Inform. Bull. Var. Stars, 5480, 1 (2003) New elements for 80 eclipsing binaries.

J/other/IBVS/5493/ table1.dat table2.dat (20Kb) Nelson R.H. IAU Inform. Bull. Var. Stars, 5493, 1 (2004) CCD minima for selected eclipsing binaries in 2003.

J/other/IBVS/5495/ notes.dat refs.dat table1.dat (19Kb) Otero S.A., Claus F. IAU Inform. Bull. Var. Stars, 5495, 1 (2004) New elements for 80 eclipsing binaries. II.

J/other/IBVS/5502/ table1.dat table2.dat (13Kb) Dvorak S.W. IAU Inform. Bull. Var. Stars, 5502, 1 (2004) Times of minima for neglected eclipsing binaries in 2003.

J/other/IBVS/5526/ table1.dat (16Kb) Price A., Vanmunster T., Starkey D., Boyd D., Zissell R., Gary B., Graham K., Macdonald II W., Aquino B., West D., Blackwell J., Walker G., Simonsen M., Henden A., Templeton M.R., Mattei J.A. IAU Inform. Bull. Var. Stars, 5526, 1 (2004) Flickering and periodic activity in the 2004 outburst of BZ UMa.

J/other/IBVS/5542/ table1.dat table2.dat table3.dat (39Kb) Dvorak S.W. IAU Inform. Bull. Var. Stars, 5542, 1 (2004) Updated elements for southern eclipsing binaries.

J/other/IBVS/5543/ position.dat table1.dat (60Kb) Diethelm R. IAU Inform. Bull. Var. Stars, 5543, 1 (2004)

163. List of minima timings of eclipsing binaries by BBSAG observers.

J/other/IBVS/5586/ notes.dat table1.dat (16Kb)

Otero S.A., Wils P., Dubovsky P.A. IAU Inform. Bull. Var. Stars, 5586, 1 (2005) New elements for 80 eclipsing binaries V.

J/other/IBVS/5592/ table.dat (10Kb) Krajci T. IAU Inform. Bull. Var. Stars, 5592, 1 (2005) Photoelectric minima of some eclipsing binary stars.

J/other/IBVS/5633/ table.dat (13Kb) Roslund C., Larsson A., Schalander P., Radbo M. IAU Inform. Bull. Var. Stars, 5633, 1 (2005) Periods of 54 known Miras and of 16 new ones in Scorpius.

J/other/IBVS/5721/ refs.dat table2.dat table3.dat table4.dat (512Kb) Kazarovets E.V., Samus N.N., Durlevich O.V., Kireeva N.N., Pastukhova E.N. IAU Inform. Bull. Var. Stars, 5721, 1 (2006) The 78th name-list of variable stars.

J/other/IBVS/5731/ eb.dat notes.dat observ.dat puls.dat table1.dat table2.dat (190Kb) Hubscher J., Paschke A., Walter F. IAU Inform. Bull. Var. Stars, 5731, 1-31 (2006) Photoelectric minima of selected eclipsing binaries and maxima of pulsating stars.

J/other/IBVS/5767/ stars.dat table1.dat (37Kb) Le Borgne J.F., Klotz A., Boer M. IAU Inform. Bull. Var. Stars, 5767, 1-7 (2007) The GEOS RR Lyr survey.

Author Index

A. J.: J/A+A/413/233 A.: J/A+A/356/529 J/A+A/364/391 J/A+A/364/391 J/A+A/364/391 J/A+A/385/1014 J/A+A/390/707

J/A+A/390/707 J/A+A/421/937 J/A+A/430/421 J/A+A/445/939 J/A+A/445/939 J/A+A/448/1075 J/A+A/475/915 J/A+AS/120/257 J/A+AS/135/145 J/A+AS/31/439 J/MNRAS/326/959

A. R.: J/ApJS/122/81 Aalders J.W.G.: II/97 Aaronson M.: J/AJ/90/2221 VII/75 Aars C.E.: J/AJ/127/3213 J/AJ/130/23 Aarum Ulvas V.: J/A+A/402/1033 Aarum V.: J/A+A/370/974 Abad C.: J/A+A/416/811 ABAD C.: J/A+AS/111/229 Abad C.: J/A+AS/136/307 J/AJ/127/1158 Abazajian K. et al.: II/259 Abbey A.F.: J/A+A/448/L9 Abe F.: J/MNRAS/330/137 Abe R.: J/PASJ/53/971 Abel T.: J/A+A/459/423 Abell G.O.: VII/110A VII/4A Abelleira P.: J/A+A/416/811

Abergel A.: J/A+A/350/883 J/A+A/357/219 J/A+A/372/173 J/A+A/420/945 J/A+A/421/623

J/A+A/469/575 Abgrall H.: J/A+A/445/361 J/A+AS/141/297 Abia C.: J/A+AS/102/361 Ables H.D.: J/AJ/103/638 J/AJ/105/1571 Abo Elazm M.S.: J/other/Ap+SS/203.53 Abraham R.: J/ApJ/499/112 J/MNRAS/297/885 Abraham R.G.: J/AJ/112/359 J/AJ/127/2455 J/ApJ/471/694 J/ApJ/551/111 J/ApJ/588/218 J/ApJS/102/289

J/ApJS/107/1 J/ApJS/113/1 J/ApJS/116/211 J/ApJS/116/231 J/ApJS/116/247 Abraham Z.: J/A+A/409/857 J/A+A/461/949 J/AJ/128/2364 J/AJ/131/951 Abrahamian et al.: II/223 Abrahamyan G.V.: III/246 J/other/Ap/46.46 Abrahamyan H.V.: J/other/Ap/44.328 J/other/Ap/44.88 J/other/Ap/45.322 J/other/Ap/46.475 Abt H.A.: III/106 ABT H.A.: III/4 Abt H.A.: J/ApJ/521/682 J/ApJ/541/841 J/ApJ/573/359 J/ApJ/616/562 J/ApJ/629/507 J/ApJS/155/175

J/ApJS/162/207 J/ApJS/28/1 J/ApJS/59/95 J/ApJS/99/135 Abuter R.: J/ApJ/643/1011 Acebo Y.: J/AJ/127/1811 Aceituno F.: J/A+A/468/L13 Achren J.: J/A+A/427/769 Achterberg A.: J/A+A/313/1008 J/other/APh/26.282 Acke B.: J/A+A/463/243 Acker A.: J/A+A/307/215 J/A+A/327/191 J/A+A/403/659 J/A+A/463/265 J/A+AS/106/559 J/A+AS/125/289

J/A+AS/132/13 V/100 V/100 V/127 V/53A V/84 ACKER A.: VI/14 Acker A.: VII/24 Ackermann G.: II/260 Ackermann M.: J/other/APh/26.282 Acosta-Pulido J.A.: J/A+A/476/L17 Actis E.L.: J/A+A/430/327 Actis E.L: J/A+AS/118/1 Actis E.L.: J/A+AS/136/1

Adami  C.:  J/A+A/381/420 J/A+A/393/439 J/A+A/403/857 J/A+A/428/1043 J/A+A/451/1159

J/A+A/457/79 J/A+A/459/679 J/A+A/461/81 J/A+A/470/161 J/A+AS/129/399 J/AJ/120/1 J/ApJS/126/209 J/MNRAS/382/279

Adams A.N.: I/100A I/100A I/100A ADAMS A.N.: I/39 Adams F.C.: J/ApJS/125/161 Adams J.: J/other/APh/26.282 J/other/APh/26.367 Adams J.D.: J/AJ/124/1570 J/ApJS/134/103 Adams M.T.: J/AJ/119/3026 J/AJ/123/1528 J/AJ/127/2228 Adams N.R.: J/ApJ/606/466 Adams W.S.: J/ApJ/81/187 Adamson A.J.: J/MNRAS/359/2 J/MNRAS/373/781 Adelberger K.: J/ApJ/626/L5 J/MNRAS/356/568 Adelberger K.L.: J/ApJ/480/568 J/ApJ/519/1 J/ApJ/592/728 J/ApJS/127/1 Adelman C.J.: J/A+AS/106/333 ADELMAN C.J.: VI/11 Adelman C.J.: VI/29A Adelman S.J.: III/89 J/A+A/329/579 J/A+A/333/952 J/A+A/335/973 J/A+A/354/899 J/A+A/357/548

J/A+A/367/297 J/A+A/367/597 J/A+A/367/859 J/A+A/368/225 J/A+A/371/1078 J/A+A/390/1023 J/A+A/394/187 J/A+A/400/1043 J/A+A/401/357 J/A+A/406/975 J/A+A/406/987 J/A+A/413/285 J/A+A/429/317 J/A+A/435/1099 J/A+A/447/685 J/A+AS/103/1 J/A+AS/106/333 J/A+AS/107/353 J/A+AS/111/41 J/A+AS/114/253 J/A+AS/122/249 J/A+AS/123/445 J/A+AS/125/219 J/A+AS/125/497 J/A+AS/125/65 J/A+AS/128/245 J/A+AS/129/563 J/A+AS/132/93 J/A+AS/133/197 J/A+AS/134/53 J/A+AS/136/111 J/A+AS/136/379 J/A+AS/137/227 J/A+AS/146/13 J/BaltA/10/589 J/BaltA/14/41 J/MNRAS/316/514 J/PASP/109/9

ADELMAN S.J.: VI/11 Adelman S.J.: VI/29A VI/34 Adelman-McCarthy J.K. et al.: II/267 II/276 Adorf H.-M.: J/AJ/112/1335 Adraou A.: J/A+A/376/1123 A.E.: J/A+A/409/541 AERTS C.: J/A+A/286/136 AErts C.: J/A+A/310/164 J/A+A/361/245 AErts C: J/A+A/367/236 AErts C.: J/A+A/393/965 J/A+A/415/241 J/A+A/440/249 J/A+A/446/583 J/A+A/449/305 J/A+A/455/259

J/A+A/457/651 J/A+A/463/243 J/A+A/475/1159 Afanasiev V.L.: J/AZh/82/420 J/MNRAS/372/1259 J/other/AISAO/32.31 Affer L.: J/A+A/461/977 Afonso C.: J/A+A/389/149 J/A+A/454/185 Afonso J.: J/AJ/131/1216 J/ApJ/624/135 J/ApJS/155/1 Afonso-Luis A.: J/AJ/129/1198 J/MNRAS/351/1290 Afsar M.: J/A+A/420/595 Aftreth O.: J/AJ/120/2331 A.G.: J/AZh/72/291 Agerer F.: J/other/IBVS/5296

Aggarwal K.M.: J/A+A/399/799 J/A+A/407/769 J/A+A/410/349 J/A+A/416/383 J/A+A/418/371 J/A+A/420/783 J/A+A/427/763 J/A+A/429/1117 J/A+A/432/1151 J/A+A/436/1141 J/A+A/439/1215 J/A+A/441/831 J/A+A/450/1249 J/A+A/460/331 J/A+A/460/959 J/A+A/463/399 J/A+A/473/995 J/ApJS/121/265 J/ApJS/123/311

Aggarwal M.: J/A+A/475/393 A.G.G.M.: J/A+A/431/523 Agostinho R.J.: J/PASP/110/1434 Agudo I.: J/A+A/476/L17 Agudo-Merida L.: J/A+A/431/779 Agueros M.A.: J/AJ/126/2209 Aguerri J.A.L.: J/A+A/471/17 J/AJ/122/38 J/ApJ/602/664 J/MNRAS/355/82 Aguilar L.: J/PASP/97/932 AGUILAR L.A.: J/AJ/107/2240 Aguilera C.: J/AJ/130/2453 J/AJ/131/1648 Aguirre A.: J/A+AS/126/73 J/ApJ/596/768 Aguirre J.: J/ApJ/638/293 Aharonian F.: J/A+A/390/39 J/A+A/406/L9 Aharonian F.A.: J/A+A/410/813 Ahrens J.: J/other/APh/26.282 Ahumada A.V.: J/A+A/418/979 J/A+A/421/991 J/A+A/440/111 J/MNRAS/335/233 J/MNRAS/340/1249

J/MNRAS/346/390 J/MNRAS/349/641 J/MNRAS/367/599 J/MNRAS/377/1737 J/other/NewA/11.262 J/PASP/117/22

Ahumada J.: J/A+A/314/448 J/A+AS/109/375 J/PASP/108/338 Ahumada J.A.: J/A+A/463/789 J/AN/326/3 J/other/RMxAA/39.39 Ahumada: J/A+A/409/541 Aiello S.: III/130 Aigrain S.: J/MNRAS/370/954 J/MNRAS/377/741 Aikawa M.: J/A+A/441/1195 Aitken: J/A+A/385/1014 A.J.: J/MNRAS/326/959 Ajhar E.A.: J/AJ/108/2087 J/AJ/129/2138 Ajiki M.: J/PASJ/57/165 J/PASJ/57/881 Ak H.: J/A+A/376/158 J/A+A/403/675 J/A+A/408/611 J/A+A/419/607 J/A+A/463/233 J/A+A/464/263

J/AN/323/31 A.K.: J/A+A/374/504 Ak S.: J/other/NewA/12.446 Ak T.: J/other/NewA/12.446 AKALIN A.: J/A+AS/106/373 Akalin A.: J/A+AS/125/407 AKAN C.: J/A+A/287/817 Akan C.: J/A+A/336/920 Akan M .C.: J/A+A/374/1056 AKAN M.C.: J/A+A/281/811 Ake T.B.: J/AJ/116/941 J/ApJS/119/83 J/ApJS/146/1 Akerlof C.: J/AJ/119/1901 Akerlof C.W.: J/AJ/127/2436 Akhperjanian A.: J/A+A/390/39 J/A+A/406/L9 Akhperjanian A.G.: J/A+A/410/813 Akhunov T. Burkhonov O.: J/A+A/440/53

Akiyama M.: J/ApJ/518/656 J/ApJ/610/128 J/ApJ/634/861 J/ApJS/148/275 J/MNRAS/372/741

J/PASJ/57/165 Akiyama S.: J/ApJ/600/716 Akras S.: J/A+A/456/599 Akujor C.E.: J/A+AS/80/215 Akylas A.: J/A+A/459/693 J/A+A/466/823 J/MNRAS/352/91 J/MNRAS/366/171 al.: II/113 Alakoz A.V.: J/AZh/79/610 Alard C.: II/243 J/A+A/349/L69 J/A+A/357/871 J/A+A/362/215 J/A+A/389/149 ALARD C.: J/A+AS/109/447 Alard C.: J/A+AS/120/275 J/A+AS/121/407 J/ApJ/552/289 Albacete Colombo J.F.: J/A+A/464/211 J/A+A/474/495 J/MNRAS/355/1237

Albayrak B.: J/A+A/364/237 J/A+A/376/158 J/A+A/400/1043 J/A+A/415/283 J/A+A/420/1039

J/other/PASA/23.154 Alberdi A.: J/A+A/365/110 J/MNRAS/379/275 Albert J-N.: J/A+A/389/149 Albert J.N.: J/A+A/454/185 Albert L.: J/AJ/119/2780 Albert S.: J/ApJ/499/517 J/ApJ/500/1059 Albinson J.S.: J/A+A/412/405 Albrecht M.: J/A+A/376/1123 Albrecht P.: J/ApJS/115/185 Albrecht R.: J/A+A/287/769 J/A+A/288/413 Albrow M.D.: J/ApJ/559/1060 J/MNRAS/349/833 J/MNRAS/362/1167 Alcaino G.: J/A+A/400/917 J/A+A/407/919 J/A+A/425/509 J/A+AS/109/479 J/A+AS/109/487 ALCAINO G.: J/A+AS/112/439 Alcaino G.: J/A+AS/119/191 J/A+AS/125/1 ALCAINO G.: J/A+AS/84/59 J/AJ/107/230 Alcaino G.: J/AJ/112/2004 J/AJ/112/2020 J/AJ/114/189 J/AJ/116/2415 J/AJ/99/1831 J/PAZh/21/596

J/PAZh/22/269 J/PAZh/22/764 J/PAZh/23/454 Alcaino: J/A+A/469/529

Alcala J.M.: J/A+A/312/439 J/A+A/319/184 J/A+A/320/185 J/A+A/353/1044 J/A+A/375/130 J/A+A/428/339 J/A+A/433/151 J/A+A/470/281 J/A+AS/114/109 J/A+AS/123/329 J/A+AS/124/449 J/A+AS/146/323

Alcock C.: J/AJ/114/326 J/AJ/127/334 J/ApJ/542/257 J/ApJ/552/289 J/ApJ/562/337 J/ApJ/598/597 Aldcroft T.: J/ApJ/644/829 Aldcroft T.L.: J/ApJ/618/123 J/ApJS/134/35 J/ApJS/140/143 J/ApJS/143/257 J/ApJS/150/165 Aldering G.: II/254 J/A+A/437/789 J/AJ/116/1094 VII/204 Alef W.: J/A+A/476/L17 Aleksandrovich N.L.: J/PAZh/26/355 J/PAZh/28/323 Alekseeva A.A.: J/PAZh/32/672 Alekseeva G.A.: III/201

Alencar S.H.P.: J/A+A/409/1037 J/A+A/409/169 J/A+A/440/595 J/A+A/463/1017 J/A+AS/125/471

J/A+AS/135/555 J/AJ/126/2971 Alessi B.: J/A+A/447/921 Alessi B.S.: B/ocl J/A+A/376/441 J/A+A/388/168 J/A+A/446/949 Alexander D.M.: J/AJ/122/2810 J/AJ/124/2351 J/AJ/126/539 J/AJ/126/632 J/AJ/127/3075 J/ApJ/640/167

J/ApJ/640/603 J/ApJS/161/21 J/MNRAS/351/1290 J/MNRAS/372/1621 Alexander J.B.: II/73 Alexander P.: J/MNRAS/376/1251 Alexander R.D.: J/A+A/401/613 Alexander T.: J/ApJ/643/1011 Alexov A.: J/AJ/126/2644 Alfaro E.J.: J/A+A/394/97 J/A+A/467/1397 ALFARO E.J.: J/AJ/106/1547 Alfaro E.J.: J/AJ/116/1801 Alha L.: J/A+A/350/513 Ali B.: J/AJ/109/709 Alikakos J.: J/A+A/456/599 Alissandrakis C.E.: J/A+A/274/555 ALKSNE Z.: III/140 III/73 ALKSNIS A.: III/140 Alksnis A.: III/227 ALKSNIS A.: III/73 Alksnis A.: J/A+AS/118/105 Alksnis O.: J/A+AS/136/453 Allain S.: J/A+A/335/183 J/A+A/349/619 Allakhverdiev A.: J/PAZh/26/843 Allam S.: J/A+AS/117/39 Allam S.S.: J/AJ/127/1811 J/AJ/127/1883 J/AJ/127/2210 J/AJ/129/2062 J/AN/321/101 J/ApJS/130/237 ALLAMANDOLA L.J.: J/ApJS/86/713 Allan A.: J/MNRAS/361/1091 Allan D.J.: J/MNRAS/274/1165 Allard F.: J/A+A/337/403 J/A+A/338/1066 J/A+A/442/281 J/AJ/126/1526 Allen D.M.: J/A+A/454/895 J/A+A/454/917 Allen: J/A+A/356/529 Allen L.: J/AJ/126/1916 J/AJ/131/1574 J/ApJ/638/897 Allen L.E.: J/AJ/109/1379 ALLEN L.E.: J/ApJ/411/188 Allen L.E.: J/ApJ/566/993 Allen M.G.: J/A+A/424/545 J/A+A/472/805 Allen P.: J/AJ/126/3007 J/AJ/128/463 J/MNRAS/372/425 Allen P.D.: J/AJ/130/1516 VII/240 Allen P.R.: J/AJ/132/891 J/AJ/133/439 Allen S.W.: J/MNRAS/274/75 J/MNRAS/301/881 J/MNRAS/306/857 J/MNRAS/318/333 Allen W.H.: J/A+A/465/331 Allende Prieto C.: J/A+A/352/555 J/A+A/420/183 J/A+A/423/1109 J/A+A/434/1125 J/A+AS/129/41

J/A+AS/131/431 J/AJ/126/1362 J/AJ/130/1097 J/ApJ/527/879 J/ApJ/562/528 Allende Prieto: J/A+A/465/271 Aller H.D.: J/A+A/402/151 J/A+A/469/451 J/AJ/130/2473 Aller L. H.: J/PASP/105/1279 J/PASP/106/745 Aller L.H.: J/A+A/419/1123 J/A+AS/139/135 Aller M.C.: J/AJ/129/2138 Aller M.F.: J/A+A/402/151 J/A+A/469/451 J/AJ/130/2473 Allington-Smith J.: J/MNRAS/273/157 Alloin D. Berlind P.: J/A+A/369/57 Alloin D.: III/219 ALLOIN D.: J/A+A/283/805 Alloin D.: J/MNRAS/382/279 Allsman R.: J/ApJ/542/257 Allsman R.A.: J/AJ/114/326 J/ApJ/552/289 J/ApJ/562/337 Allsopp J.: J/MNRAS/379/663 ALLWRIGHT J.W.B.: J/ApJS/73/671 Allwright J.W.B.: J/PASP/99/739 Almaini O.: J/AJ/129/2074 J/MNRAS/343/293 J/MNRAS/351/1290 J/MNRAS/355/97 J/MNRAS/372/1621 Almeida C.: J/ApJ/624/135 J/ApJS/155/1 Almoznino E.: J/ApJS/111/143 J/MNRAS/313/641 J/MNRAS/324/580 J/MNRAS/332/441 J/MNRAS/357/645 Aloisi A.: J/ApJS/144/21 ALONSO A.: J/A+AS/107/365 Alonso A.: J/A+AS/117/227 J/A+AS/131/209 Alonso E.: J/A+A/430/327 J/A+AS/118/1 J/A+AS/136/1 Alonso Medina A.: J/A+A/422/1109 Alonso M.V.: J/AJ/108/1987 J/AJ/118/1131 J/AJ/125/2307 J/AJ/126/2268 J/AJ/127/2114 Alonso O.: J/A+AS/118/7 J/A+AS/120/323 J/ApJS/105/343 J/ApJS/122/415 Alonso R.: J/A+A/468/637 Alonso-Herrero A.: J/AJ/119/2274 J/ApJ/640/167 Alonso-herrero A.: J/ApJ/642/126 J/ApJS/151/387 Alonso-Herrero A.: J/ApJS/154/155 J/MNRAS/338/508 J/MNRAS/338/525 Alpar M.A.: J/MNRAS/269/849 Alsac L.: J/A+A/430/373 Alshino A.: J/MNRAS/382/279 Al-Shukri A.M.: J/AJ/111/393 Altamore A.: J/A+A/317/712 J/A+A/374/1085 J/A+A/420/571 J/A+AS/92/685 ALTAMORE A.: J/ApJS/71/983 Altavilla G.: J/MNRAS/369/1880 Alten V.: J/ApJ/478/603 ALTENHOFF W.J.: J/A+A/281/161 Altenhoff W.J.: J/A+AS/35/23 Althaus L.G.: J/A+A/465/249 Altieri B.: J/A+A/357/839 J/A+A/369/57 J/A+A/407/791 J/A+A/436/457 J/MNRAS/382/279 Altmann A.: J/A+A/442/1023 Altmann M.: J/A+A/357/471 J/A+A/400/877 J/A+A/461/509 J/ApJS/115/185 J/ApJS/162/1 Altschuler D.R.: J/A+AS/65/267 J/AJ/113/905 ALTSCHULER D.R.: VIII/26 Altschuler D.R.: VIII/57 Altunin V.: J/PASJ/52/997 Alvarado F.: J/A+A/400/917 J/A+A/407/919 J/A+A/425/509 J/A+A/469/529 ALVARADO F.: J/A+AS/112/439 Alvarado F.: J/A+AS/119/191 J/A+AS/125/1 ALVARADO F.: J/AJ/107/230 Alvarado F.: J/AJ/112/2004 J/AJ/112/2020 J/AJ/114/189 J/AJ/116/2415 J/AJ/99/1831 J/PAZh/21/596

J/PAZh/22/269 J/PAZh/22/764 Alvarez C.: J/A+A/416/811 J/AJ/127/1181 J/ApJS/155/123 Alvarez C.V.: J/ApJS/155/149 Alvarez H.: J/A+A/327/325 J/AJ/89/1897 Alvarez M.: J/A+A/426/247 J/A+A/468/637 Alvarez R.: J/A+A/317/761 J/A+A/374/968 J/A+A/379/305 J/A+AS/140/55 Alves D.: J/A+A/423/97 J/AJ/114/326 Alves D.R.: J/AJ/121/318 J/AJ/127/334 J/ApJ/542/257 J/ApJ/552/289 J/ApJ/562/337 J/ApJ/598/597 Alves J.: J/A+A/441/195 J/A+A/462/L17 J/A+A/466/137 J/A+A/476/199 J/AJ/111/1964 J/ApJ/573/366 Alves J.F.: J/A+A/453/101 J/AJ/128/1254 J/AJ/131/527 Alves-Brito A.: J/A+A/435/657 Alvey N.D.S.: J/A+A/410/597 Alzner A.: J/A+AS/132/237 J/A+AS/132/253 Amado P.J.: J/A+A/450/715 Amadon A.: J/A+A/389/149 Amanullah R.: J/A+A/437/789 Amati L.: J/A+A/455/813 Ambrocio-Cruz P.: J/AJ/124/3157 Amico P.: J/A+AS/120/257 J/A+AS/131/105 J/AJ/127/3121 Amieux G.: III/192 J/A+AS/110/107 Amirkhanyan V.R.: VIII/49 VIII/49 VIII/49 Ammar P.: J/AJ/130/1977 Ammons S.M.: J/ApJ/637/1102 J/ApJS/169/430 Amnuel P.R.: IX/8 Amosse A.: J/A+A/410/343 Amram P.: J/A+A/330/881 J/A+A/420/147 J/A+A/455/107 J/A+A/465/1099 AMRAM P.: J/A+AS/113/35 Amram P.: J/A+AS/137/419 Amrose S.: J/AJ/119/1901 A.N.: J/AZh/77/420 An J.H.: J/MNRAS/351/1071 Anandarao B.G.: J/AJ/126/2963 Ananjevskaja J.K.: J/A+A/396/125 J/A+A/401/531 J/A+A/451/901 Ananthamurthy S.: J/other/BASI/32.295 Ananyevskaya J.K.: II/123A V/52 and Broderick J.J.: J/AJ/102/1663 and Corbin T.E.: I/231 and Feibelman W. A.: J/PASP/106/745 and Forman W.: J/ApJ/511/65 and Ohishi N.: J/PASJ/59/245 and Siebert J.: J/A+A/319/413 Anderlucci E.: J/A+AS/121/119 Andernach H. et al.: VIII/43 Andernach H.: J/A+A/336/35 J/A+A/370/468 J/A+AS/123/119 J/AJ/122/2222 J/AJ/123/51 J/MNRAS/269/301

J/MNRAS/311/456 J/other/BSAO/49.53 VII/165A VIII/50 Andernach: VIII/43 Anders P.: J/A+A/401/1063 Andersen A.C.: J/A+A/423/983 Andersen J. Astron. Astrophys. 384: J/A+A/384/491 ANDERSEN J.: III/102 III/103 III/104 III/105 III/127 III/127 III/134 III/79 Andersen J.: J/A+A/275/101 ANDERSEN J.: J/A+A/291/121 J/A+A/299/53 Andersen J.: J/A+A/304/415 J/A+A/319/481 J/A+A/322/460 J/A+A/324/91 J/A+A/375/30 J/A+A/384/491

J/A+A/386/187 J/A+A/389/149 J/A+A/403/1105 J/A+A/409/1037 J/A+A/416/1117 J/A+A/423/189 J/A+A/454/185 J/A+A/469/285 J/A+A/473/829 J/A+AS/102/603 J/A+AS/118/407 J/A+AS/125/471 J/A+AS/126/21 J/A+AS/128/251 J/A+AS/130/245 J/AJ/114/1087

Andersen J.; Nordstroem B. ; Loden et al.: III/40A Andersen J.: V/117 Andersen M.I.: J/A+A/304/415 J/A+A/322/460 J/A+A/374/1049 J/A+AS/118/407 J/A+AS/140/135 Andersen T.B.: J/A+AS/53/181 Anderson B.: J/MNRAS/358/1159 Anderson C.: J/ApJ/644/829 Anderson C.M.: VI/105 Anderson E.: III/116 J/AJ/109/2318 Anderson J.: J/A+A/444/831 J/ApJ/595/187 Anderson J.E.: J/AJ/121/2331 Anderson J.E.Jr: J/AJ/121/2308 Anderson J.E.JR: J/AJ/121/31 Anderson K.S.J.: J/AJ/116/1626 J/AJ/127/2085 Anderson M.: J/other/SerAJ/166.13Short Anderson M.C.: J/ApJ/441/307 J/ApJ/456/234 Anderson N.: J/ApJS/60/507 Anderson S.F.: J/AJ/121/2308 J/AJ/122/503 J/AJ/123/3409 J/AJ/123/567 J/AJ/124/1651 J/AJ/126/1023

J/AJ/126/1720 J/AJ/126/2125 J/AJ/126/2209 J/AJ/126/2579 J/AJ/127/2838 J/AJ/129/2542 J/ApJ/450/51 J/ApJ/607/426 J/ApJS/155/257 J/ApJS/165/1

Andersson B.-G.: J/AJ/124/2164 Ando H.: J/A+A/468/663 J/PASJ/57/109 J/PASJ/57/13 J/PASJ/57/165 Andrade M.: J/AJ/127/1181 Andre P.: J/A+A/350/883 J/A+A/359/113 J/A+A/372/173 J/A+A/372/L41 J/A+A/401/613 J/A+A/420/945

J/A+A/421/623 J/A+A/458/789 J/A+A/472/519 Andre Philippe M.: J/A+A/357/219 Andreani P.: J/A+A/321/123 ANDREANI P.: J/A+AS/112/347 Andreani P.: J/A+AS/136/471 Andreasyan R.R.: J/other/Ap/43.413 Andreazza C.M.: J/A+AS/116/21 Andrei A. H.: J/A+A/425/1107

Andrei A.H.: J/A+A/384/650 J/A+AS/117/335 J/A+AS/135/227 J/A+AS/136/531 J/A+AS/143/265

J/AJ/125/2728 Andreon S.: J/A+A/323/337 J/A+A/467/73 J/A+AS/116/429 J/A+AS/126/67 J/A+AS/141/113

J/MNRAS/382/279 Andreuzzi G.: J/A+A/337/17 J/A+A/380/478 J/A+A/390/961 J/A+A/425/509 J/AJ/121/916 Andrews P.: II/266 J/AJ/127/1158 Andrews S.M.: J/AJ/128/1233 ANDREZ R.: I/74 ANDRIEVSKY S.: J/A+A/293/347 Andrievsky S.M.: J/A+A/281/465 J/A+A/305/551 J/A+A/381/32 J/A+A/392/491 J/A+A/413/159

J/AJ/128/343 J/AJ/129/433 J/AJ/130/1880 Andrillat Y.: III/183 III/217 J/A+AS/117/281 J/A+AS/118/495 J/A+AS/120/99 J/A+AS/124/441

J/A+AS/128/475 J/A+AS/131/469 J/A+AS/131/479 Andronova A.A.: I/266 Andruchow I.: J/A+A/409/857 Andruk V.: J/AN/316/225 Andruk V.N.: J/other/KFNT/22.231 Angeli C.A.: J/A+AS/134/463 Angelini L.: IX/12 IX/31 J/A+A/448/L9 Angeretti L.: J/MNRAS/358/397 Anglada G.: J/ApJ/463/205 J/ApJ/628/789 J/ApJS/125/427 ANGUITA C.: I/43 Anguita C.: I/82 J/AJ/112/2408 Ankay A.: J/PAZh/26/843 Ann H.B.: J/other/NewA/11.293 Annear P.R.: II/221A Annibali F.: J/A+A/433/497 J/A+A/445/79 J/A+A/463/455 J/MNRAS/364/146

Annis J.: J/AJ/121/2308 J/AJ/121/2331 J/AJ/121/31 J/AJ/123/1807 J/AJ/123/567 J/AJ/125/1817

J/AJ/126/2209 J/AJ/127/1811 J/AJ/127/1883 J/AJ/130/968 J/ApJS/148/243 Annuk K.: J/A+A/282/137 J/A+A/415/273 Ansari R.: J/A+A/303/137 J/A+A/389/149 J/A+A/421/509 J/A+A/454/185 ANSARI R.: J/A+AS/109/447 Ansari S.G.: J/A+AS/109/267 Antezana R.: J/AJ/112/2408 ANTHONY-TWAROG B.J.: J/AJ/107/1577

Anthony-Twarog B.J.: J/AJ/109/2828 J/AJ/116/1922 J/AJ/119/2282 J/AJ/125/1383 J/AJ/127/1000

J/AJ/129/872 J/AJ/131/461 ANTHONY-TWAROG B.J.: J/PASP/106/486 Anthony-Twarog B.J.: J/PASP/107/32 Antia B.: J/AJ/107/751 Antico G.: J/A+AS/127/445 Antilogus P.: J/A+A/437/789 Antipin S.V.: III/229

Antipova L.I.: J/AZh/72/855 J/AZh/72/864 J/AZh/73/862 J/AZh/75/586 J/AZh/77/96 J/AZh/78/807

J/AZh/79/909 J/AZh/80/704 J/AZh/81/658 J/AZh/82/601 Antokhin I.: J/AJ/109/817 Antokhin I.I.: J/A+A/418/841 J/A+A/477/593 Antokhina E.A.: J/PAZh/25/764 Anton S.: J/MNRAS/329/877 Antonelli A.: J/A+A/472/705 J/ApJ/570/100 Antonelli L.A.: J/A+A/455/813 J/A+A/469/503 J/A+A/474/793 J/MNRAS/327/771

Antonello E.: J/A+A/314/541 J/A+A/333/L35 J/A+A/363/29 J/A+A/367/759 J/A+A/385/793

J/A+A/445/901 ANTOV A.: J/A+AS/103/183 Antov A.: J/A+AS/116/1 Anupama G.C.: J/MNRAS/372/1315 Aoki K.: J/ApJS/89/57 J/PASJ/57/165 Aoki W.: J/A+A/328/175 J/A+A/353/528 J/ApJ/638/1018 J/PASJ/57/13 J/PASJ/58/389 A.P.: J/A+A/364/391 Apai D.: J/A+A/434/987

Aparicio A.: J/A+A/240/262 J/A+A/333/479 J/A+A/417/597 J/A+A/432/851 J/A+AS/144/5

J/A+AS/145/451 J/A+AS/71/297 APARICIO A.: J/AJ/106/1547 Aparicio A.: J/AJ/110/212 APPELQUIST L.: III/79 Appen-Schnur G.F.O.v.: J/A+AS/119/153 Appenzeller I.: J/A+A/323/707 J/A+A/398/49 J/A+A/418/885 J/A+A/421/41 J/ApJS/117/319 Appleton P.N.: J/AJ/131/250 Appodia B.: J/A+AS/125/459 Appourchaux T.: J/A+A/478/197 Apriamashivili S.P.: III/236 Aquino B.: J/other/IBVS/5526 A.R.: B/cgps Arabadjis J.S.: J/ApJ/613/1179 Aracil B.: II/243 Aragon-Salamanca A.: J/A+A/379/798 J/A+A/427/397 J/A+A/444/365 J/ApJ/528/123 J/MNRAS/311/456

J/MNRAS/338/508 J/MNRAS/338/525 J/MNRAS/361/109 J/MNRAS/366/144 J/MNRAS/371/1777 Araki H.: J/PASJ/52/631 Arany-Prado L.I.: J/A+A/440/1153 Araujo F.X.: J/A+A/477/877 Araya E.: J/ApJS/154/553 Arbutina B.: J/other/PASA/23.154 Arce H.G.: J/AJ/131/527 J/ApJ/517/264 Archer S.: VIII/1A Archinal B.A.: J/AJ/110/880 ARDEBERG A.: III/105 Ardeberg A.: III/69A J/A+A/465/357 J/A+AS/30/261 J/A+AS/73/37 Ardila D.R.: J/AJ/128/1990 J/ApJ/612/202 J/ApJ/621/53 J/ApJS/150/1 Arditti F.: J/A+AS/122/489 Aref’ev V.A.: J/PAZh/28/323 Arellano Ferro A.: J/A+A/368/250 J/A+A/407/1059 J/A+A/408/L29 J/A+A/471/515 J/other/RMxAA/42.75 Arena R.: J/A+AS/109/201 Arenou F.: I/196 J/A+A/345/471 J/A+AS/137/451 Arens J.: II/77 Arentoft T.: J/A+A/374/1056 J/A+A/389/931 J/A+A/418/249 J/A+A/438/201 J/A+A/465/965 Aretxaga I.: J/MNRAS/372/1621 Argelander F.W.A.: I/122 Argiroffi C.: J/A+A/439/1149 J/A+A/459/199 Argon A.L.: J/ApJ/615/702 J/ApJS/160/220 ARGUE A.N.: I/169 Argue A.N.: I/196 J/MNRAS/259/563 ARGUE A.N.: VI/1 ARGYLE R.W.: I/129 Argyle R.W.: I/303 J/A+A/312/1031 J/A+A/469/1221 J/AJ/130/2701 J/MNRAS/367/1170 Arias E.F.: J/AJ/127/3587 Arias J.: J/AJ/120/2101 Arias J.I.: J/MNRAS/366/739

Arimoto N.: J/A+A/394/927 J/A+A/459/423 J/ApJ/527/573 J/ApJ/528/123 J/MNRAS/365/1392

J/MNRAS/366/144 J/PASJ/56/295 Aristidi E.: J/A+A/367/865 J/A+AS/134/545 J/ApJS/131/561 J/MNRAS/367/1170 Arkharov A.A.: III/201 J/A+A/462/1081 J/PAZh/30/243 J/PAZh/32/628 J/PAZh/32/672 ARKHIPOVA V.: VII/100 Arkhipova V.: VII/62A

Arkhipova V.P.: II/229 J/AZh/78/963 J/PAZh/25/30 J/PAZh/26/705 J/PAZh/27/187 J/PAZh/27/841 J/PAZh/29/858 J/PAZh/30/297 J/PAZh/30/855 J/PAZh/32/48 J/PAZh/32/662 J/PAZh/32/737 J/PAZh/33/679 VII/236

Arlot J.-E.: J/A+A/314/312 J/A+A/380/727 J/A+A/383/719 J/A+A/391/767 J/A+A/400/1095 Arlot J.E.: J/A+A/427/371 J/A+AS/125/399 ARMANDROFF T.E.: J/AJ/106/986 Armandroff T.E.: J/AJ/110/2131 J/AJ/110/2715 J/AJ/121/841 J/AJ/123/2511 J/AJ/124/1464 J/AJ/124/601

J/AJ/127/318 J/AJ/129/2232 Armond T.: J/AJ/127/1117 Armsdorfer B.: J/A+A/413/1037 Armus L.: J/AJ/121/97 J/AJ/131/250 J/ApJ/633/871 J/ApJ/634/128 J/ApJ/634/L1

Arnaboldi M.: J/A+A/386/140 J/A+A/403/857 J/A+A/428/1043 J/A+A/428/339 J/A+A/439/413

J/A+A/457/79 J/MNRAS/369/120 Arnal E.M.: J/AJ/121/413 VIII/75 VIII/76 Arnal: J/A+AS/142/25 Arnaud K.A.: III/153 J/ApJS/96/303 J/MNRAS/356/568 Arning C.A.: J/AJ/115/2587 Arnold R.: J/A+A/445/503 Arnould M.: J/A+A/441/1195 Arnouts S.: J/A+A/341/641 J/A+A/379/436 J/A+A/379/740 J/A+A/384/81 J/A+A/392/741 J/A+A/428/1043

J/A+A/456/881 J/A+A/457/79 J/A+A/462/553 J/AJ/120/2206 J/AJ/122/2190 ARNTH-JENSEN N.: J/AJ/106/56 Arosio I.: J/A+A/446/839 Arp H.: J/A+A/319/33 Arp H.C.: VII/13 VII/170 ARP H.C.: VII/74A Arquilla R.: J/ApJS/99/223 Arriagada P.: J/A+A/476/73 ARRIBAS S.: J/A+AS/107/365 Arribas S.: J/A+AS/117/227 J/A+AS/131/209 Arsenijevic J.: J/A+AS/125/399 Artigau E.: J/A+A/456/905 Artiukhina N.M.: II/140 Arts J.N.: J/MNRAS/346/627 Artyukh V.S.: J/A+A/403/555 J/A+A/421/455 VIII/67 Artyukhina N.M.: II/205 Arzner K.: J/A+A/468/353 Arzoumanian Z.: J/ApJ/615/897 A.S.: J/A+A/444/895 Ascenso J.: J/A+A/466/137 J/A+A/476/199 Aschenbach B.: IX/10A IX/29 Asensio Ramos A.: J/A+A/465/339 Ashimbaeva N.T.: III/182 Ashman K.M.: J/AJ/118/752 J/ApJ/513/733 Ashok N.M.: J/AJ/116/1447 J/AJ/117/2980 Ashoka B.N.: J/A+A/455/303 Asiain R.: J/A+AS/115/401 Aslan Z.: J/BaltA/11/465 ASONUMA A.: VII/140 Aspelund-Johansson J.: J/A+A/435/373 Aspin C.: J/A+AS/106/165 Asplund M.: J/A+A/353/287 J/A+A/415/993 J/A+A/423/1109 J/A+A/440/321 Asplund M. Lambert D.L.: J/A+A/390/235 Assafin M.: J/A+A/446/949 J/A+AS/117/335 J/A+AS/136/531 J/AJ/125/2728 Assef R.: J/ApJS/162/1 Assemat F.: J/A+A/435/507 ASSENDORP R.: J/A+AS/100/473 Assendorp R.: J/A+AS/117/39 Astier N.: J/A+AS/31/439 Astier P.: J/A+A/437/789 Astronomical Data Analysis Center: B/subaru Astrophysical Research Consortium (ARC): III/241 Atanackovic-Vukmanovic O.: J/other/PASA/23.154 ATHANASE M.: I/74 Athanassoula E.: J/A+A/415/849 J/A+A/421/595 J/A+A/464/L21 Athreya R.: J/ApJS/107/19 Athreya R.M.: J/A+A/463/519 Atlee D.W.: J/other/APh/26.282 Atteia J.-L.: J/other/GCN/1070.1 Attolini M.R.: J/MNRAS/276/57 Attridge J.: V/68A Atwell J.: J/AJ/129/872 Aubourg E.: J/A+A/303/137 J/A+A/389/149 J/A+A/454/185 AUBOURG E.: J/A+AS/109/447 Audard M.: J/A+A/468/353 J/A+A/468/379 Audinos P.: J/A+A/314/871 Aufdenberg J.P.: J/A+A/453/155 J/ApJ/484/394 J/ApJ/636/1002 AUGARDE R.: J/A+A/285/1 Augusteijn T.: J/A+A/312/93 J/A+A/466/1031 J/A+AS/126/401 J/MNRAS/362/1006 Augusto P.: J/MNRAS/368/1411 Auman J.R.: VII/213 Aurass H.: J/A+AS/119/489 Auriere M.: J/A+A/421/509 J/A+A/475/1053 J/A+AS/46/347 Aussel H.: J/A+A/342/313 J/A+A/440/61 J/A+A/454/453 J/A+AS/141/257 J/ApJ/602/571 J/MNRAS/325/1173

J/MNRAS/358/397 J/MNRAS/382/279 Ausseloos M.: J/A+A/455/259 J/A+A/463/243 Authors Publ. Astron. Soc. Pac. 110: J/PASP/110/1342 Authors Publ. Astron. Soc. Pac. 111: J/PASP/111/221 Authors (various journals) PASAu: J/other/PASAu/5.290 A.V.: IX/36 J/A+A/409/541 Avedisova V.S: V/112 Avedisova V.S.: V/74 Avera E.: J/AJ/114/122 Avila R.: J/A+AS/134/545 Avila-Reese V.: J/A+A/412/633 J/AJ/129/682 Aviles R.: J/AJ/112/2408 Avni Y.: J/ApJS/92/53 Avruch I.M.: J/PASJ/52/997 AWADALLA N.S.: J/A+A/289/137 Axelrod T.: J/ApJ/542/257 Axelrod T.S.: J/AJ/114/326 J/AJ/127/334 J/ApJ/552/289 J/ApJ/562/337 Axelsson M.: J/A+A/452/975 AXER M.: J/A+A/300/751 AXON D.J.: II/178 Axon D.J.: J/MNRAS/358/1159 Aydin C.: J/A+A/400/1043 Ayvazyan V.T.: J/A+AS/147/169 J/AJ/124/1283 Azzaro M.: J/ApJ/645/228 J/ApJS/141/61

Azzopardi M.: J/A+A/377/945 J/A+AS/102/251 J/A+AS/102/451 J/A+AS/97/603 J/other/Ap/41.356 J/other/Ap/44.328 J/other/Ap/44.88 J/other/Ap/45.322 J/other/Ap/46.475 V/13

B. Cram: J/MNRAS/296/839 B.: J/A+A/403/247 J/A+A/404/423 J/A+AS/126/319 B. Vollmer: VIII/74A BAADE D.: J/A+A/294/135 Baade D.: J/A+A/459/137 Baade R.: J/A+A/391/L11 Baan W.: J/A+A/424/371 Baas F.: J/A+AS/101/513 BABA A.: VII/140 Baba D.: J/ApJ/596/1064 J/MNRAS/370/1979 Baba N.: J/other/PNAOJ/2.561 Babadzhanyants M.K.: J/A+A/356/L21 J/AZh/72/447 J/AZh/79/675 J/PAZh/25/893 Babbedge T.: J/AJ/129/1198 J/AJ/129/2074 J/MNRAS/351/1290 J/MNRAS/358/397 J/MNRAS/372/1621 Babel J.: J/A+A/394/151 Babenko Yu: J/A+A/438/377 Babler B.: VI/105 Babler B.L.: J/ApJ/635/560 Babu G.S.D.: J/other/BASI/32.295 Babu J.G.: J/ApJ/508/314 Babusiaux C.: J/A+A/442/365 Bach U.: J/A+A/402/151 J/A+A/476/L17 Bacher A.: J/A+A/413/1037 Bachev R.: J/ApJS/145/199 Bachilla R.: J/AJ/127/1360 Bachiller R.: J/A+A/319/235 J/ApJS/160/272 J/AZh/71/37 Backer D. C.: J/ApJS/121/171 Backman D.: J/PASP/118/1690 Bade N.: J/A+A/310/384 J/A+A/358/77 J/A+A/401/927 BADE N.: J/A+AS/110/469 Bade N.: J/A+AS/127/145 Badiali M.: J/A+A/374/1085 Badnell N.R.: J/A+A/430/331 J/A+A/466/763 J/A+A/469/1203 Baes M.: J/MNRAS/352/478 J/MNRAS/357/819 Baessgen G.: I/207 Baffa C.: J/A+A/327/755 J/A+AS/120/521 J/A+AS/137/101 J/A+AS/142/65 J/AJ/112/2786 Baganoff F.: J/AJ/112/2408 Baganoff F.K.: J/ApJ/589/225 J/ApJ/613/1179 Baggett S.M.: J/AJ/116/1626 J/AJ/127/2085 Baggett W.E.: J/AJ/116/1626 J/AJ/127/2085

Baggley G.: J/AJ/112/2408 J/ApJS/106/1 J/MNRAS/292/499 J/MNRAS/305/259 J/MNRAS/326/473

J/MNRAS/327/249 Bagildinsky B.K.: I/242 Bagnulo S.: J/A+A/374/265 J/A+A/450/763 J/A+A/475/1053 J/MNRAS/367/1478 J/MNRAS/376/361 Bagnuolo W.G.: J/AJ/106/352 J/ApJ/525/420 Bagnuolo W.G.JR: J/ApJ/639/1069 Bahcall J.N.: J/AJ/107/1745 BAHCALL J.N.: J/AJ/108/1786 Bahcall J.N.: J/AJ/113/669 J/AJ/116/1757 J/ApJ/409/28 J/other/APh/26.282 J/PASP/106/646 VII/149 Bahcall N.: J/AJ/116/2081 Bahcall N.A.: J/AJ/121/2308 J/AJ/121/2331 J/AJ/121/31 J/AJ/122/503 J/AJ/123/1807 J/AJ/123/567

J/AJ/125/1817 J/AJ/126/2209 J/AJ/126/2579 J/AJ/127/1811 J/AJ/130/968 J/AJ/131/1 J/AJ/131/49 J/ApJ/409/28 J/ApJS/148/243 J/ApJS/163/80 VII/149

Bai G.S.: J/A+A/425/671 Bai J.M: J/A+AS/132/83 J/A+AS/136/455 Bai J.M.: J/ApJ/522/846 Baier F.W.: VII/89B VII/89B BAIESI-PILLASTRINI G.C.: VII/70A Baiesi-Pillastrini G.C.: VII/71 Bailer-Jones C.A.L.: J/A+A/396/513 J/A+A/417/557 J/A+A/430/1005 Baillon P.: J/A+A/421/509 J/MNRAS/351/1071 Bailyn C.: J/AJ/112/2408 J/AJ/127/1158 Bailyn C.D.: J/ApJ/608/L93 Bains I.: J/MNRAS/367/1609 Bair X.: J/other/APh/26.282 Baird S.: J/AJ/119/1824 Bajaja E.: VIII/75 VIII/76 Bajaja: J/A+AS/142/25 Bajkova A.T.: J/AZh/83/821 Baker A.C.: J/MNRAS/351/1290 Baker A.J.: J/A+A/406/593 Baker R.I.: J/ApJS/154/673 Bakis V.: J/A+A/417/263 Bakker E.J.: J/A+A/306/924 J/A+A/307/869 J/A+A/310/893 J/A+A/323/469 J/A+A/336/263 Bakos G.: J/AJ/128/1761 Bakos G.A.: I/279 J/AcA/50/221 J/AJ/127/861 J/AJ/129/2856 J/MNRAS/372/1657 Balachandran S.: J/AJ/106/1059 J/AJ/110/1229 Balachandran S.C.: J/A+A/372/862 Balaguer-Nunez L.: J/A+A/381/464 J/A+A/426/819 J/A+A/426/827 J/A+A/437/457 J/A+A/470/585 Balaguer-Nunnez L.: J/A+AS/133/387 Balam D.D.: J/A+A/466/1031 Balayan S.K.: J/A+A/414/487 Balazs B.: VII/101A BALAZS B.: VII/31B VII/44B Balcells M.: J/A+A/390/863 BALCELLS M.: J/AJ/107/135 Balcells M.: J/ApJS/154/155 Bald A.: J/A+A/466/31 Baldacci L.: J/A+A/431/1189 Baldi A.: J/A+A/409/79 J/MNRAS/320/387 Baldry I.K.: J/MNRAS/352/939 Baldwin J.A.: J/AJ/112/2408 J/ApJ/644/1006 J/ApJS/149/157 Baldwin J.E.: J/MNRAS/246/110 BALDWIN J.E.: VIII/18 Baldwin J.E.: VIII/1A BALDWIN J.E.: VIII/21 VIII/22 VIII/23 VIII/24 VIII/25 Baldwin R.J.: II/68 Baldwin T.J.: II/68 Bale K.: J/A+A/475/1053 Balega I.: J/A+A/422/627 Balega I.I.: J/A+A/385/87 J/A+AS/105/503 J/A+AS/140/287 Balega Yu.: J/A+A/448/1075

Balega Y.Y.: J/A+A/366/868 J/A+A/385/87 J/A+A/409/245 J/A+A/422/627 J/A+AS/105/503

J/A+AS/140/287 J/AJ/127/1181 Balin N.: J/A+A/379/162 Baliunas S.L.: J/A+A/325/1115 J/AJ/111/439 BALIUNAS S.L.: J/ApJ/416/787 Baliunas S.L.: J/PASP/96/707 Balklavs A.: III/227 Balkowski C.: J/A+A/297/617 J/A+A/330/881 J/A+A/342/1 J/A+A/405/99 J/A+A/408/465 J/A+A/408/67

J/A+A/415/9 J/A+A/455/107 J/A+A/464/L21 J/A+A/465/1099 J/A+AS/100/583 BALKOWSKI C.: J/A+AS/113/35 Balkowski C.: J/A+AS/130/323 J/ApJ/520/491 BALKOWSKI C.: VII/159 Ballereau D.: J/A+A/378/861 Ballester G. E.: J/PASP/108/682 Ballester G.E.: J/AJ/113/1001 J/AJ/117/1700 Bally J.: J/A+A/470/605 J/AJ/127/1117 J/AJ/130/1763 J/ApJ/478/603 J/ApJ/575/354 J/ApJS/160/319

J/ApJS/160/511 VIII/10 VIII/28 Balog Z.: J/A+A/463/1017 J/AJ/131/527 Balogh M.: J/MNRAS/337/256 Balogh M.L.: J/A+A/421/59 J/ApJ/549/820 J/MNRAS/318/703 J/MNRAS/335/10 J/MNRAS/360/587 Balona L.: J/AJ/120/2101 BALONA L.A.: J/A+A/294/135 Balona L.A.: J/A+A/309/787 BALONA L.A.: J/MNRAS/260/782 BALONEK T.J.: J/AJ/107/1453 Balonek T.J.: J/AJ/116/2805 Balsano R.: J/AJ/119/1901 J/AJ/127/2436 Balser D.S.: J/ApJS/165/338 Balsley J.: J/PASP/111/685 Baltay C.: J/AJ/127/1158 Balucinska-Church M.: J/MNRAS/344/134 VI/62 BALUTEAU J.-P.: J/A+A/303/175 Baluteau J.-P.: J/A+A/406/L47 Balzano V.A.: J/ApJS/62/751 Bamba A.: J/ApJ/635/214 Bamford S.P.: J/MNRAS/361/109 J/MNRAS/366/144 Band D.L.: J/ApJS/126/19 Bandiera R.: J/other/PASA/19.455 Bandyopadhyay R.M.: J/ApJS/165/173 Bange J.F.: J/A+A/406/751 Bania T.M.: J/ApJ/517/282 J/ApJ/582/756 J/ApJ/635/560 J/ApJS/165/338 Banks G.D.: J/AJ/123/873 J/AJ/124/690 J/MNRAS/346/787 J/MNRAS/355/1303 BANZI M.: J/A+AS/103/245 Bappu M.K.V.: II/40 Baptista R.: J/A+AS/125/399 Baraffe I.: J/A+A/327/1039 J/A+A/337/403 Barannikov A.A.: J/PAZh/25/209 Barat C.: IX/36 Baratta G.B.: J/A+A/258/423 J/A+AS/92/685 BARATTA G.B.: J/ApJS/71/983 Baratta G.B.: VI/48 Barazza F.D.: J/A+A/388/29 J/A+A/445/765 Barba R.: J/AJ/120/2101 Barba R.H.: J/AJ/125/1940 J/MNRAS/355/1237 J/MNRAS/366/739 Barban C.: J/A+A/405/1095 J/A+A/463/243 J/A+A/478/197 Barber C.R.: J/MNRAS/274/1165 Barber R.J.: VI/119 Barber T.: J/MNRAS/377/787 Barbera C.: J/A+A/415/849 J/A+A/421/595 Barbera M.: J/ApJ/414/846 Barbier D.: III/187 Barbier M.: III/38 Barbier-Brossat M.: III/108 III/118 III/160A III/161 BARBIER-BROSSAT M.: III/184 Barbier-Brossat M.: III/213 Barbieri C.: J/A+A/287/769 J/A+A/288/413 BARBIERI C.: J/A+AS/112/347 Barbieri C.: VII/3 VII/69 Barbieri M.: J/A+A/384/879 J/A+A/404/187 J/A+A/420/683 Barblan F.: J/A+A/398/1073 J/A+AS/132/367 BARBON R.: II/103 Barbon R.: II/159 II/227 J/A+A/317/423 J/A+A/415/145 J/A+AS/115/325 BARBUY B.: J/A+A/286/444 J/A+A/296/680 J/A+A/297/503 J/A+A/300/726

Barbuy B.: J/A+A/304/347 J/A+A/306/134 J/A+A/308/733 J/A+A/319/850 J/A+A/326/614 J/A+A/331/70 J/A+A/364/674 J/A+A/381/32 J/A+A/390/931 J/A+A/391/179 J/A+A/397/177 J/A+A/400/533 J/A+A/403/1105 J/A+A/404/223 J/A+A/404/661 J/A+A/416/1117 J/A+A/429/607 J/A+A/435/657 J/A+A/450/105 J/A+A/454/895 J/A+A/454/917 J/A+A/465/799 J/A+AS/106/161

BARBUY B.: J/A+AS/108/653 J/A+AS/111/17

Barbuy B.: J/A+AS/120/153 J/A+AS/122/483 J/A+AS/127/471 J/A+AS/145/399 J/AJ/130/55

J/ApJS/124/527 J/MNRAS/289/745 J/MNRAS/338/253 VI/120 Barbuy: J/A+AS/126/319 Barcons X.: J/A+A/428/383 J/A+A/444/79 J/A+A/469/27 J/A+A/476/1191 J/ApJ/498/77 J/ApJ/559/654

J/MNRAS/311/456 Bardelli S.: J/A+A/337/17 J/A+A/379/35 J/A+A/385/39 J/A+A/403/857 J/A+A/419/71 J/A+A/428/1043

J/A+A/442/841 J/A+A/457/79 J/A+A/463/519 J/AN/316/39 J/MNRAS/285/898 J/MNRAS/296/599 J/MNRAS/312/540 J/MNRAS/320/387 J/MNRAS/324/1131

Barden M.: J/MNRAS/351/1290 Barden S.C.: J/PASP/111/1233 Barentine J.C.: J/A+A/444/643 J/AJ/131/1674 J/AJ/132/676 Baret B.: J/other/APh/26.282 Bareyre P.: J/A+A/303/137 J/A+A/389/149 J/A+A/454/185 BAREYRE P.: J/A+AS/109/447 Barger A.J.: J/AJ/122/2810 J/AJ/126/539 J/AJ/126/632 J/AJ/127/180 J/AJ/127/3121 J/AJ/127/3137

J/AJ/128/1483 J/AJ/128/1501 J/ApJ/497/188 J/ApJ/554/742 Barker M.K.: J/AJ/128/224 Barkhouse W.A.: J/AJ/121/2843 J/ApJ/618/123 J/ApJ/644/829 J/ApJ/645/955 Barklem P.S.: J/A+A/420/183 J/A+A/435/373 J/A+A/439/129 J/A+AS/142/467 Barlow D.J.: J/AJ/129/1001 Barlow E.J.: J/ApJ/607/L33 J/ApJ/636/765 J/ApJS/170/175

Barlow M.J.: J/A+A/382/184 J/A+A/442/249 J/A+AS/146/437 J/ApJ/497/330 J/MNRAS/234/583

J/MNRAS/368/1959 Barlow T.A.: J/AJ/130/825 J/ApJ/619/L95 J/ApJS/164/38 J/PASP/113/10 Barman T.: J/A+A/471/L51 J/A+A/475/1125

Barmby P.: J/A+A/403/493 J/AJ/115/6 J/AJ/119/727 J/AJ/122/2458 J/AJ/131/1416 J/AJ/131/527

J/ApJ/610/247 J/ApJ/640/167 J/ApJ/642/126 J/ApJS/154/155 J/MNRAS/371/221 Barnard E.E.: VII/220A Barnbaum C.: J/A+A/314/871 J/ApJ/518/859 J/ApJS/105/419 Barneoud R.: J/AJ/108/405 Barnes D.G.: J/AJ/123/873 J/AJ/124/1954 J/AJ/124/690 J/MNRAS/337/641 Barnes III T.G.: J/A+A/440/487 J/ApJS/156/227 Barnes J.: J/MNRAS/379/816 Barnes J.R.: J/MNRAS/379/647 Barnes J.V.: III/116 Barnes S.: J/AJ/112/1054 J/AJ/113/1045 J/AJ/122/1486 Barnes T. G. III: J/PASP/111/812 Barnes T.G. III: J/ApJS/117/135 Barnes T.G.: J/A+A/330/515 J/A+A/476/73 J/A+AS/125/313 J/A+AS/125/321 Barney I.: I/141 Barney Y. et al.: I/3 Baron N.: J/A+A/391/767 Barone F.: J/A+A/318/797 Barr J.: J/AJ/130/445 Barr P.: J/A+A/357/839 J/A+A/369/57 VI/43 BARRADO D.: J/A+AS/111/275 Barrado y Navascues D.: J/A+A/310/879 J/A+A/326/780 J/A+A/395/813 J/A+A/436/853 J/A+A/471/499

J/ApJ/546/1006 Barrado Y Navascues D.: J/ApJ/610/1064 Barrado y Navascues D.: J/ApJS/134/103 Barrena R.: J/A+A/416/839 J/A+A/444/685 J/A+A/449/461 J/A+A/455/45 J/A+A/467/37 J/A+A/469/861

J/A+A/476/L17 Barrera L.: J/A+A/380/347 Barret D.: J/A+A/400/521 J/A+A/445/155 J/ApJS/154/585 Barrett P.: J/A+A/412/405 Barrett P.E.: J/other/SAAOC/15.78 Barrientos F.: J/AJ/112/2408 Barrientos L.F.: J/AJ/134/1103 J/ApJ/627/621 J/ApJS/153/397 Barrio J.A.: J/A+A/390/39 Barris B.: J/AJ/130/2453 J/AJ/131/1648 J/ApJ/594/1 Barris B.J.: J/ApJ/602/571 J/MNRAS/327/1004 Barroso Jr. J.: J/A+AS/125/399 Barry D.J.: J/AJ/106/352 J/AJ/108/2299 J/AJ/111/936 J/AJ/117/1890 J/ApJ/635/280 Barsella B.: III/130 Barsony M.: J/ApJ/630/381 J/ApJS/112/109 Barstow M.A.: J/ApJS/119/207 J/ApJS/147/145 Barsukova E.A.: J/AZh/79/309 J/AZh/84/147 J/PAZh/28/764 Bartasiute S.: J/BaltA/10/319 J/BaltA/11/321 J/BaltA/11/417 J/BaltA/11/465 J/BaltA/8/443 Bartaya R.A.: III/112 BARTAYA R.A.: III/72 Bartelmann M.: J/AJ/127/1811 Bartelt M.: J/other/APh/26.282 Barth A. J.: J/ApJ/510/659 Barth A.J.: J/AJ/121/3048 J/ApJS/115/185 Barthel P.D.: J/A+A/319/757 J/A+A/463/97 BARTHEL P.D.: J/A+AS/108/287 Barthelmy S.D.: J/ApJ/633/L77 Barthes D.: J/A+AS/140/55 Bartholdi P.: J/A+A/398/1073 J/A+AS/132/367 BARTKEVICIENE D.: V/72 Bartkeviciene V.-D.: IV/20 BARTKEVICIUS A.: II/157A

Bartkevicius A.: II/93 III/125 IV/20 J/AN/323/139 J/BaltA/10/481 J/BaltA/11/153 J/BaltA/5/1

J/BaltA/8/325 BARTKEVICIUS A.: V/72 Bartko F.: J/AJ/128/1990 J/ApJ/612/202 J/ApJ/621/53 J/ApJS/150/1 Bartlett R.E.: J/ApJ/492/62 Bartlett T.J.: II/68 Barton E.J.: J/AJ/116/1573 J/AJ/131/527 J/AJ/132/197 Barton Gillespie E.: J/ApJ/582/668 BARTUNOV O.S.: II/176 Bartunov O.S.: II/201 II/218 II/256 Bartus J.: J/A+AS/125/11 Bartzakos P.: J/MNRAS/324/33 Barucci M.A.: J/A+A/375/275 Barwick S.W.: J/other/APh/26.282 Barziv O.: J/A+A/377/925 BASCHEK B.: J/A+A/289/740 Bash F.N.: VIII/42 Basilakos S.: J/MNRAS/344/602 J/MNRAS/349/135 J/MNRAS/351/1290 BASINSKI J.: II/119 Baskill D.: J/A+A/476/1191 Basri G.: J/A+A/409/169 J/A+A/440/595 J/AJ/126/1526 J/AJ/132/663 Bassani L.: J/ApJ/607/L33 J/ApJ/636/765 J/ApJS/106/399 J/ApJS/121/473 J/ApJS/170/175 Bassino L.P.: J/A+A/355/138 J/MNRAS/367/156 Bastian N.: J/A+A/397/473 J/A+A/435/65 BASTIAN U.: I/132 Bastian U.: I/146 BASTIAN U.: I/159A Bastian U: I/193 BASTIAN U.: I/194 I/206 Bastian U.: I/207 I/208 I/210 I/250 Bastian U: I/259 BASTIAN U.: III/155 Bastian U.: III/182 Bastien P.: J/A+A/335/973 J/ApJ/604/717 Bastin R.J.: J/MNRAS/368/1959 Batalha C.: J/A+A/409/1037 J/A+A/409/169 Batalha C.C.: J/A+AS/128/561 Batirshinova V.M.: J/A+A/399/105 BATRAKOV Yu.V.: I/118 Batrakov Yu.V.: I/124 BATRAKOV Yu.V.: I/127 I/139 I/145 I/183B I/184C Batrakov Yu.V.: I/204 BATRAKOV Yu.V.: I/217 I/233 Batrakov Yu.V.: I/245 Battaglia G.: J/A+A/459/423 Battaner E.: J/A+A/382/513 J/A+A/391/519 J/A+A/399/457 Batten A.H.: B/sb9 V/122 BATTEN A.H.: V/40 Batten A.H.: V/64

Battinelli P.: J/A+A/278/36 J/A+A/357/437 J/A+A/410/795 J/A+A/416/111 J/A+A/417/479

J/A+A/418/33 J/A+A/424/125 J/A+A/436/91 J/A+A/447/473 J/A+A/456/905 J/A+A/473/143 BATTINELLI P.: J/A+AS/104/379 Battinelli P.: J/AJ/120/1801 J/AJ/123/832 J/AJ/125/1298 J/AJ/125/3037 BATTY M.J.: VIII/45 Battye R.A.: J/MNRAS/360/340 J/MNRAS/376/371 Baturina G.D.: I/44 Batuski D.: J/ApJ/416/36 J/ApJS/115/1 Batuski D.J.: J/AJ/124/1918 J/ApJ/520/491 Batyrshinova V.M.: J/PAZh/25/115 Baudrand A.: J/A+A/375/275 J/A+A/423/381 Baudry A.: J/A+A/325/255 J/A+A/394/975 Bauer A.: J/AJ/123/567 Bauer C.: J/A+AS/125/399 Bauer F.: J/A+A/389/149 J/ApJ/640/603 Bauer F.E.: J/A+A/469/27 J/AJ/122/2810 J/AJ/124/2351 J/AJ/126/539 J/AJ/126/632 J/ApJ/640/167

J/ApJS/129/547 J/ApJS/161/21 J/ApJS/165/173 Baugh C.M.: J/MNRAS/329/87 J/MNRAS/352/939 Baum S. A.: J/ApJS/122/81 Baum S.A.: J/A+A/463/97 J/A+AS/131/303 J/AJ/119/486 J/ApJ/448/521 Baum W.A.: J/AJ/105/1196 J/ApJ/456/174 J/ApJ/468/633 Baume G.: J/A+A/371/908 J/A+A/381/884 J/A+A/402/549 J/A+A/407/527 J/A+A/409/933 J/A+A/417/961

J/A+A/419/149 J/A+A/432/491 J/A+A/436/527 J/A+A/442/917 J/A+A/444/831 J/A+A/445/493 BAUME G.: J/A+AS/106/339

Baume G.: J/A+AS/116/75 J/A+AS/124/13 J/A+AS/137/233 J/MNRAS/346/18 J/MNRAS/367/1441

J/other/RMxAA/39.89 Baume G.L.: J/A+A/430/471 Baumgardt H.: J/A+A/466/931 J/A+AS/146/251 J/ApJ/602/264 Baumgart C.W.: J/AJ/91/530 Bautista M.: J/A+A/365/268 J/A+A/418/1171 Bautista M.A.: J/A+A/410/359 J/A+A/414/377 J/A+A/420/763 J/A+A/469/1203 Bautz M.: J/ApJ/381/33 Bautz M.W.: J/AJ/126/539 J/ApJ/554/742 J/ApJ/589/225 J/ApJ/613/1179 Baxter D.: J/A+A/287/769 J/A+A/288/413 Bay R.: J/other/APh/26.282 Baykal A.: J/MNRAS/269/849 Baylac M.-O.: J/A+A/393/897 Baylac M.O.: J/A+AS/137/451 Bayne G.: J/MNRAS/331/609 Bayne G.P.: J/MNRAS/349/833 Bazterra V.: J/A+AS/109/201 Bazzano A.: J/A+A/458/21 J/ApJ/607/L33 J/ApJ/636/765 J/ApJS/170/175 B.D.: V/111 Beardmore A.P.: J/A+A/448/L9 Beasley A.J.: J/ApJS/141/13 Beasley M.A.: J/AJ/131/814 J/MNRAS/345/949 J/MNRAS/355/608 Beattie K.: J/other/APh/26.282 Beatty A.E.: V/92A Beauchamp A.: J/ApJS/93/187 Beauchemin M.: J/ApJ/509/309 BEAUCHEMIN. M.: J/MNRAS/247/182 Beaulieu J.P.: J/A+A/303/137 Beaulieu J-P.: J/A+A/389/149 Beaulieu J.P.: J/A+A/454/185 BEAULIEU J.-P.: J/A+AS/109/447 Beaulieu J.-P.: J/MNRAS/371/1405 Beaulieu S.F.: J/ApJ/515/610 V/127 Beaver E.A.: J/AJ/116/941 Bec-Borsenberger A.: I/196 J/A+A/375/275 J/A+A/391/767 J/A+A/406/751 J/A+A/423/381 Bechtold J.: J/AJ/112/62 J/ApJ/457/102 J/ApJ/571/665 J/ApJS/130/37 J/ApJS/140/143 J/ApJS/91/1

J/ApJS/95/1 Beck R.: J/A+A/472/785 J/AJ/131/250 BECK S.C.: J/A+A/299/89 Beck T.L.: J/AJ/117/1375 Becka T.: J/other/APh/26.282 Becker A.: J/ApJ/542/257 J/ApJ/598/597 Becker A.C.: J/AJ/114/326 J/AJ/127/334 J/AJ/130/2453 J/AJ/131/1648 J/ApJ/552/289 J/ApJ/562/337 Becker J.K.: J/other/APh/26.282 Becker K.-H.: J/other/APh/26.282 Becker R.H.: J/A+A/356/445 J/AJ/112/407 J/AJ/113/2000 J/AJ/121/1872 J/AJ/121/31 J/AJ/127/2565

J/AJ/129/348 J/AJ/130/156 J/AJ/130/586 J/AJ/131/1 J/AJ/131/1948 J/AJ/131/666 J/ApJ/461/127 J/ApJ/505/L7 J/ApJ/607/60 J/ApJS/126/133 J/ApJS/135/227 J/ApJS/143/1 J/ApJS/74/181 J/ApJS/91/347 VIII/13 VIII/17 VIII/48 VIII/51 VIII/59 VIII/71

Becker W.: J/ApJ/615/887 J/ApJ/615/897 Beckerman E.: J/AJ/127/1959 Becklin E.E.: J/ApJ/555/932 J/ApJS/161/394 Beckman J.: J/A+A/374/394 J/A+AS/131/105 Beckman: J/A+AS/135/145 Beckman J.E.: J/A+A/354/823 J/A+A/414/23 J/A+A/430/911 J/A+A/459/L13 J/A+A/466/905 BECKMAN J.E.: J/A+AS/93/557 J/AJ/106/56 Beckman J.E.: J/ApJ/491/L31 J/ApJS/111/203 J/ApJS/145/355 Beckmann V.: J/A+A/358/77 J/A+A/401/927 J/A+A/467/585 J/A+AS/127/145 J/ApJ/635/214 Beckwith S.: J/PASP/118/1690 Beckwith S.V.W.: J/AJ/129/1534 Bedalov A.: J/A+A/468/1039 Bedding T.R.: J/A+A/410/879 Bedin L.R.: J/A+A/444/831 J/A+A/471/515 BEDIN V.N.: I/43 Bedin V.N.: I/82 Bedin V.S.: I/44 Beele D.: II/202 J/A+A/304/415 Beers T.: J/A+A/416/1117

Beers T.C.: J/A+A/397/899 J/A+A/403/1105 J/A+A/422/527 J/A+A/431/143 J/A+A/439/129 J/A+AS/117/317 J/AJ/103/1987 J/AJ/103/267 J/AJ/108/268 J/AJ/108/538 J/AJ/109/874 J/AJ/112/1188 J/AJ/117/2308 J/AJ/117/2329 J/AJ/117/981 J/AJ/119/2866 J/AJ/126/1362 J/AJ/127/1567 J/AJ/130/1097 J/AJ/130/2804 J/AJ/90/2089 J/ApJ/562/528 J/ApJS/103/433 J/ApJS/123/639 J/ApJS/76/1001 J/ApJS/96/175 J/MNRAS/320/451 J/PASP/102/1392 J/PASP/103/973

Behrend R.: J/A+A/465/331 Beichman C.A.: II/246 J/A+A/359/865 J/AJ/124/1570 J/AJ/125/2521 J/ApJ/476/311 J/ApJ/620/1010

J/ApJ/636/1098 BEICHMULLER F.: I/84 Beilicke M.: J/A+A/406/L9 J/A+A/410/813 Beintema D.: J/A+A/390/1033 BEINTEMA D.A.: II/148 Beirao P.: J/A+A/438/251 Bejar V.J.S.: J/AN/325/705 Belanger G.: J/ApJ/636/765 J/ApJS/170/175 Belczynski K.: J/A+AS/146/407 Beletsky Y.: J/A+A/466/137 Beletsky Yu.V.: J/A+A/392/491 Belik S.I.: J/A+A/411/559 J/A+A/427/933 J/AJ/129/433 J/MNRAS/371/879 Belikov A.N.: J/A+A/332/575 J/A+A/384/145 J/A+AS/134/525 V/85A Belkin I.N.: J/A+AS/105/503 Bell D.J.: J/ApJS/152/251 J/PASP/111/1169 Bell E.F.: II/253 J/ApJ/586/794 J/ApJ/633/844 Bell G.R.: J/ApJS/162/346 Bell J.F.: J/ApJS/83/87 J/MNRAS/328/17 J/MNRAS/335/275 J/MNRAS/342/1299 Bell K.L.: J/A+A/430/725 J/A+A/432/731 J/A+A/436/1131 J/A+A/452/1113 J/A+AS/125/543 Bell K.R.: V/73A BELL R.A.: II/160 Bell R.A.: II/19 J/AJ/117/247 J/AJ/119/1424 J/AJ/119/1448 J/AJ/127/1227 BELL R.A.: VI/55 Bell S.: J/MNRAS/362/1006 Bellas-Velidis I.: J/A+A/423/919 J/MNRAS/351/1290 Bellas-Velidis Y.: J/A+AS/116/1

Bellazzini M.: J/A+A/416/917 J/A+A/420/173 J/A+A/456/985 J/A+A/464/201 J/A+A/471/127 J/A+A/473/429 J/AJ/122/2569 J/AJ/124/3222 J/AJ/124/915 J/ApJ/654/915 J/MNRAS/349/1278 J/MNRAS/354/708 J/MNRAS/357/265 J/MNRAS/360/185

Belley J.: J/A+A/321/363 Belloche A.: J/A+A/472/519 BELLONI P.: J/A+A/297/61 Belloni P.: J/A+AS/118/65 Belloni T.: J/A+A/368/835 J/A+A/418/509 Belmonte J.A.: J/A+A/468/637 Beloborodov A.M.: J/ApJ/643/1011 Belokon E.T.: J/A+A/356/L21 Belokon’ E.T.: J/AZh/79/675 Belokon E.T.: J/PAZh/25/893 Belokonj E.T.: J/AZh/72/447 Belous M.L.: II/206 Belskaya I.N.: J/A+AS/113/115 Beltran M. T.: J/A+A/474/911 Beltran M.T.: J/A+A/432/921 J/A+A/447/221 Bely-Dubau F.: J/A+AS/105/77 Belzer E.: J/ApJ/598/375 BEM J.: J/A+AS/110/411 Bembrick C.: J/A+A/465/331 Ben Nessib N.: J/A+A/423/397 J/A+A/434/773 Benaglia P.: J/A+A/370/468 BENCIVENNI D.: J/AJ/102/137 Bender C.F.: J/A+A/371/79 Bender: J/A+A/395/753 Bender R.: J/A+A/395/431 J/A+A/398/49 J/A+A/407/423 J/A+A/415/123 J/A+A/418/885 J/A+A/420/97

J/A+A/421/41 J/A+A/444/365 J/A+A/445/423 J/A+A/465/375 J/A+AS/141/449 J/AJ/114/1365 J/AJ/129/2138 J/ApJ/621/673 J/MNRAS/326/473

Bendik P.: J/A+A/447/735 Bendo G.J.: J/ApJ/633/871 Benedict G.F.: III/86 J/A+A/476/73 J/AJ/123/1411 BENEST D.: J/A+A/299/621 Benetti S.: J/MNRAS/369/1880 Benevides-Soares P.: J/A+A/361/1143 J/A+A/375/308 J/A+A/383/1054 J/A+A/409/361 J/A+AS/136/261

J/A+AS/136/531 J/A+AS/139/321 Beninati S.: J/A+A/413/1177 Benitez N.: J/A+A/432/381 J/AJ/128/1990 J/ApJ/612/202 J/ApJ/621/53 J/ApJS/150/1 Benitez T.: J/ApJ/607/665 Benjamin R.A.: J/ApJ/635/560 Benko J.M.: J/AcA/50/221 J/MNRAS/372/1657 Benn C.R.: J/MNRAS/200/747 BENN C.R.: J/MNRAS/272/699 Benn C.R.: J/MNRAS/348/857 J/MNRAS/370/1034 VIII/30 Bennert N.: J/A+A/456/953 J/A+A/459/55 BENNETT A.S.: VII/10 Bennett A.S.: VIII/1A Bennett C.L.: J/ApJS/148/97 J/ApJS/51/67 VIII/52B Bennett D.: J/ApJ/542/257 J/ApJ/598/597 Bennett D.P.: J/AJ/114/326 J/AJ/127/334 J/ApJ/552/289 J/ApJ/562/337 Bennett J.: J/PASP/110/779 Bennett K.: J/A+AS/143/145 Bennett T.: J/A+A/465/331 Bennie P.J.: J/ApJS/115/185 Benoist C.: J/A+A/379/740 J/A+A/384/81 J/A+A/392/741 J/A+A/399/813 J/A+A/408/905 J/A+A/409/439

J/A+A/430/19 J/A+A/435/781 J/A+A/442/841 J/A+A/452/119 J/A+A/456/881 J/A+A/461/81 J/A+A/462/553 J/A+AS/137/75 J/A+AS/137/83

Benoni T.: J/A+A/371/943 J/A+AS/141/491 Bensby T.: J/A+A/410/527 J/A+A/415/155 J/A+A/433/185 J/A+A/467/665 Bensch F.: VIII/66 VIII/66 Benson J.: J/PASJ/52/997 Benson K.: J/A+A/472/805 BENSON P.J.: J/AJ/107/1453 Benson P.J.: V/68A BENTOLILA C.: III/96 Bentz M.C.: J/ApJ/632/799 BENTZ W.: III/105 Benz A.O.: J/A+A/475/549 Berdnikov L.N.: II/217 J/A+A/397/997 J/A+AS/143/211 J/AJ/130/1194 J/ApJS/137/209 Berdyugin A.: J/A+A/384/1050 J/A+A/454/277 Berdyugin A.V.: J/A+A/372/276

Berdyugina S.V.: J/A+A/334/863 J/A+A/374/1049 J/A+A/390/179 J/A+A/394/505 J/A+A/432/657

J/A+A/440/735 J/A+A/476/881 BERDYUGINA S.V.: J/A+AS/107/9 Berdyugina S.V.: J/AZh/72/855 BERGBUSCH P.A.: J/AJ/106/1024 Bergbusch P.A.: J/AJ/112/1061 BERGBUSH P.A.: VI/55 Bergeat J.: J/A+A/321/236 J/A+A/321/L9 J/A+A/332/L53 J/A+A/342/773 J/A+A/344/263 J/A+A/369/178

J/A+A/390/967 J/A+A/429/235 Berger D.H.: J/A+A/453/155 Berger J.: II/111 II/239 III/64 Bergeron J.: J/A+A/283/759 J/A+A/309/81 J/A+A/393/425 J/A+A/451/457 J/ApJ/595/685 J/ApJS/139/369

J/ApJS/155/271 J/ApJS/161/21 J/MNRAS/379/1343 Bergeron L.E.: J/PASP/107/846 J/PASP/109/998 Bergeron P.: J/A+A/459/557 J/ApJ/631/1100 J/ApJS/133/413 J/ApJS/156/47 Berghaus P.: J/other/APh/26.282 Berghoefer T.: J/A+AS/99/591 J/MNRAS/306/279 Berghoefer T.W.: J/A+A/384/890 J/A+AS/118/481 Bergman P.: J/A+A/358/257 J/A+A/473/177 J/A+AS/144/441 J/ApJS/117/427 Bergmann D.J.: J/A+A/371/79 Bergmann M.: J/AJ/130/445 Bergner Yu.K.: J/A+AS/147/5

Bergond G.: J/A+A/448/155 J/A+A/449/937 J/A+A/462/507 J/A+A/464/L21 J/A+A/470/505

J/A+A/472/121 Bergstroem E.J.: J/ApJ/549/441 Bergvall N.: J/A+A/335/85 J/A+AS/137/419 Berkhuijsen E. M.: II/144 II/145 Berley D.: J/other/APh/26.282 Berlin A.B.: J/A+AS/139/545 J/AZh/62/229 J/AZh/84/227 Berlind A.A.: J/ApJS/143/73 Berlind P.: J/AJ/129/856 J/AJ/129/907 J/AJ/131/527 J/ApJ/510/659 J/ApJ/626/L5 J/PASP/111/438 Bernabei S.: J/A+A/399/271 J/A+A/459/557 Bernabeu G.: J/A+A/372/477 J/A+AS/119/271 Bernacca P.L.: III/30A Bernard A.: J/A+AS/123/507 Bernard J.P.: J/A+A/469/575 Bernard J.-P.: J/ApJS/95/419 Bernardi M.: J/AJ/121/2308 J/AJ/122/503 J/AJ/123/567 J/AJ/125/1817 J/AJ/125/2307 J/AJ/126/2268

J/AJ/130/968 J/AJ/131/1288 J/AJ/131/2018 J/PASJ/55/771 Bernardini E.: J/other/APh/26.282 Bernard-Salas: J/A+A/431/523 Bernasconi L.: J/A+A/465/331 Bernasconi P.A.: J/A+A/325/167 J/A+AS/128/471 J/A+AS/135/405 Bernazzani M.: J/A+AS/61/473 Berne O.: J/A+A/469/575 Bernhard K.: J/MNRAS/368/1757 Bernloehr K.: J/A+A/390/39 J/A+A/406/L9 J/A+A/410/813 Bernstein G.M.: J/ApJS/96/117 Bernstein R.: J/AJ/131/2114 Bernstein R.A.: J/ApJ/571/56 Berriman G.: VII/75 Berrington K.A.: J/A+A/373/987 J/A+A/422/731 J/A+A/430/331 J/A+AS/109/193 Berrington R. C.: J/AJ/115/708 BERRINGTON R.C.: J/AJ/107/1868 Berrington R.C.: J/AJ/123/2261 Bersanelli M.: J/A+A/397/213 J/A+A/410/847 J/AJ/104/28 Bershady M.A.: J/A+A/469/1211 J/AJ/126/1183 J/ApJ/505/50 J/ApJS/109/45 Bersier D.: J/A+A/309/446 Bersier D: J/A+A/341/818 Bersier D.: J/A+A/381/32 J/A+A/453/309 J/A+A/476/73 BERSIER D.: J/A+AS/108/25 J/A+AS/108/9 Bersier D.: J/AJ/123/840 J/MNRAS/371/1405 Bersier: J/A+A/428/587 Berta S.: J/A+A/434/149 J/A+A/440/61 J/A+A/457/501 J/A+A/467/565 J/AJ/129/2074 Bertaud Ch.: III/46 Bertaux J.-L.: J/A+A/474/293

Bertelli G.: J/A+A/240/262 J/A+A/329/943 J/A+A/392/1129 J/A+AS/106/275 J/A+AS/141/371

J/AJ/110/212 VI/96 Bertero M.: J/AJ/129/1534 Berthier J.: J/A+A/465/331 J/A+AS/125/399 Berthomieu G.: J/A+A/478/197 BERTIAU F.C.: III/2 Bertin E.: J/A+A/439/413 Bertin G.: J/A+AS/120/257 Bertola F.: J/A+A/342/671 J/A+A/360/439 J/A+A/374/394 J/A+A/382/488 J/A+A/408/873 J/A+A/416/507

J/A+A/424/447 J/A+A/440/107 J/A+AS/120/257 J/A+AS/129/583 J/A+AS/131/105 Bertoldi F.: J/A+A/435/575 J/A+AS/129/87 J/AJ/128/1974 J/ApJ/466/254 J/ApJ/606/664 Bertout C.: J/A+AS/129/337 Bertram D.: J/MNRAS/274/1165 J/MNRAS/274/1194 Bertram R.: J/A+A/369/57 J/ApJ/510/659 J/ApJS/115/185 Bertrand D.: J/other/APh/26.282 Bertrand J.-F.: J/AJ/109/817 Bertsch D.L.: J/ApJS/101/259 J/ApJS/107/227 J/ApJS/123/79 J/ApJS/157/324 J/ApJS/94/551 Bertschinger E.: J/ApJS/106/1 J/MNRAS/292/499 Berulis I.I.: J/AZh/71/37 Beser N.: II/230 Beshenov G.V.: J/AZh/80/8 Besier S.C.: J/MNRAS/331/609

Bessell M.S.: J/A+A/333/231 J/A+A/439/129 J/A+A/446/515 J/A+AS/134/489 J/AJ/114/2644 J/AJ/115/734 J/AJ/119/1748 J/AJ/120/333 J/AJ/123/290 J/AJ/124/2039 J/AJ/127/1014 J/AJ/128/1684 J/MNRAS/306/361 J/MNRAS/310/982 J/MNRAS/311/741

Besson D.: J/other/APh/26.367 Besson D.Z.: J/other/APh/26.282 Bessonova L.A.: V/42 Best P.: J/A+A/362/519 J/A+A/444/365 J/A+A/456/791 J/ApJ/591/640 J/MNRAS/371/1777 Best P.N.: J/A+A/471/1105 J/MNRAS/346/627 J/MNRAS/362/9 J/MNRAS/366/1265 J/MNRAS/372/1621 Bestman W.: J/AJ/108/268 Bestul S.M.: I/100A BESTUL S.M.: I/39 Beswick R.: J/A+A/472/805 Betlem H.: J/A+AS/128/179 Betrix F.: J/A+A/324/949 Bettens R.P.A.: J/A+AS/119/111 Bettoni D.: J/A+A/387/26 J/A+A/393/389 J/A+A/405/5 J/A+A/420/873 J/A+A/422/941 J/A+A/454/453

J/A+A/470/39 J/A+AS/124/61 J/ApJ/542/673 Betzler A.S.: J/A+AS/134/463 Beuermann K.: J/A+A/335/467 J/A+A/347/47 J/A+A/350/805 J/A+A/458/541 J/A+A/460/783 Beuing J.: J/A+A/395/431 Beurle K.: J/A+AS/121/65 J/A+AS/136/257 Beuther H.: J/A+A/466/1065 J/A+A/466/989 J/A+A/474/911 J/A+A/475/925 Beuzit J.-L.: J/A+A/364/665 Beuzit J.L.: J/A+A/375/L27 J/A+A/410/1039 J/A+A/410/1051 J/A+A/414/351 Beyer P.J.: J/AJ/130/23 Bhat N.D.R.: J/ApJ/605/759 Bhathal R.: J/AJ/123/873 J/AJ/124/1954 J/AJ/124/690 J/AJ/128/16 J/MNRAS/337/641 Bhatia A.K.: J/A+A/444/305 Bhatt B.C.: J/A+AS/122/111 J/MNRAS/358/1290 J/other/BASI/35.53 Bhatt H.C.: J/A+A/402/963 BHATT H.C.: J/A+AS/111/399 Bhatt H.C.: J/MNRAS/355/1272 J/MNRAS/358/1290 J/MNRAS/369/1822 J/MNRAS/379/1237 Bhavsar S.P.: J/AJ/118/2014 Bian Y.: VII/168 Bianchi L.: J/A+A/366/498 J/AJ/121/2020 J/AJ/130/825 J/AJ/132/378 J/ApJ/469/629 J/ApJ/619/L95

J/ApJ/633/871 J/ApJS/122/431 J/ApJS/164/38 Bianchi S.: J/AJ/122/2190 Biazzo K.: J/A+A/454/301 J/AN/328/938 Bibo E.A.: J/A+AS/121/223 Bica E.: III/219 BICA E.: J/A+A/283/805 J/A+A/286/444 J/A+A/296/680 J/A+A/300/726 Bica E.: J/A+A/306/134 J/A+A/308/733 J/A+A/319/850 J/A+A/326/614 J/A+A/331/70 J/A+A/359/347

J/A+A/360/529 J/A+A/383/631 J/A+A/390/931 J/A+A/391/179 J/A+A/404/223 J/A+A/429/607 J/A+A/435/657 J/A+A/440/111 J/A+A/450/105 J/A+AS/106/161

BICA E.: J/A+AS/108/653 J/A+AS/111/17

Bica E.: J/A+AS/120/153 J/A+AS/122/483 J/A+AS/126/319 J/A+AS/127/423 J/A+AS/127/471

J/A+AS/145/399 J/AJ/116/723 J/AJ/119/1214 J/ApJS/116/263 J/MNRAS/289/745 J/MNRAS/335/233 Bica E.L.D.: J/AJ/117/238 J/ApJS/101/41 J/MNRAS/338/253 Bica: J/A+A/397/177 J/A+A/400/533 Bicay M. D.: J/ApJS/125/349 Bicay M.D.: J/A+AS/147/169 J/ApJS/98/369 Bicker J.: J/A+A/413/37 BIDELMAN W.P.: III/119 Bidelman W.P.: III/46 J/ApJS/112/557 BIDELMAN W.P.: VI/15 VI/32 BIEHL D.: J/A+AS/42/391 BIEMONT E.: J/A+AS/102/435 J/A+AS/113/71 Bien R.: I/210 Bienayme O.: J/A+A/398/141 J/AZh/84/429 J/AZh/86/158 J/MNRAS/371/879 BIERMANN P.L.: J/A+A/298/375 Biermann P.L.: J/other/APh/26.282 Biggs A.D.: J/MNRAS/371/963 VIII/72 BIGGS E.S.: III/4 Bignall H.: J/MNRAS/323/757 Bignami G.F.: J/A+A/458/245 Bijaoui A.: J/A+AS/49/715 BIJAOUI A.: VII/159 Bik A.: J/A+A/397/473 Bikmaev I.F.: J/A+AS/136/453 J/AZh/81/541 Bildsten L.: J/ApJ/578/405 Bilir S.: J/AN/326/321 J/MNRAS/349/1069 J/MNRAS/355/307 J/MNRAS/357/497 J/MNRAS/360/1345

J/MNRAS/373/1483 J/other/NewA/12.446 Bills T.E.: J/A+A/371/79 Biman J.M.: J/MNRAS/379/1237 Binello A.M.: J/A+A/370/1071 J/A+AS/131/153 Binggeli B.: J/A+A/388/29 J/A+A/448/983 J/AJ/90/1681 Bird A.J.: J/ApJ/607/L33 J/ApJ/636/765 J/ApJS/170/175 Bird C.M.: J/AJ/109/874 Biretta J. A.: J/ApJS/122/81

Birkinshaw M.: J/A+A/456/791 J/A+A/476/1191 J/AJ/98/1148 J/ApJ/591/640 J/MNRAS/326/473

J/MNRAS/382/279 Birkle K: J/A+A/389/439 Birkle K.: J/A+AS/109/537 Birkmann S.M.: J/A+A/474/883 Birmingham A.: J/AJ/123/1528 Biro I B: J/A+A/403/637 Biro I.B.: J/A+A/349/515 J/A+A/394/97 Biryukov V.V.: III/209 Bischof O.B.: J/AJ/113/2000 Bischoff K.: J/ApJ/510/659 BISCHOFF M.: I/83 IV/5 Bischoff M.: IV/6 IV/9 BISCHOFF M.: V/78 Bisnovatyi-Kogan G.S.: J/A+AS/114/341 Bisterzo S.: J/ApJ/645/613 Biviano A.: J/A+A/310/31 J/A+A/310/8 J/A+A/407/791 J/A+A/416/839 J/A+A/442/841 J/A+A/444/685

J/A+A/451/1159 J/A+A/461/397 J/A+A/470/39 J/A+AS/111/265 J/A+AS/122/409 J/A+AS/129/399 Bixler J.V.: J/A+A/250/370 Bizau J.-M.: J/A+A/439/387 Bizyaev D.: J/A+A/389/795 Bizzocchi L.: J/A+A/425/767 J/A+A/455/1161 Bjorkman K.S.: J/A+A/397/1035 J/A+A/408/305 J/A+AS/147/5 J/ApJ/598/L43 Bjornsson C.-I.: J/A+A/385/55 Blaauw A.: J/AJ/117/354 Black J.H.: J/A+A/432/369 J/AJ/112/62 Blacker B.: J/AJ/112/1335 Blackwell D.E.: J/A+A/282/899 J/A+AS/129/505 Blackwell J.: J/other/IBVS/5526 Blackwell-Whitehead R.: J/A+A/416/655 Blades J. C.: J/ApJS/122/355 Blades J.C.: III/157 J/A+A/287/769 J/A+A/288/413 Blagoev K.: J/A+A/388/704 Blagrave K.P.M.: J/ApJ/644/1006 Blain A.: J/MNRAS/344/887 J/MNRAS/372/1621 Blain A.W.: J/ApJ/614/671 J/ApJ/616/71 J/ApJ/622/772 J/MNRAS/345/1 Blair W.P.: J/ApJS/146/1 Blais K.: V/68A Blais-Ouellette S.: J/A+A/420/147 Blake C.: J/A+A/374/1056 J/A+A/399/365 J/MNRAS/358/30 J/MNRAS/375/68 Blake G.A.: J/ApJ/606/L73 J/ApJS/158/188 Blake M.: J/A+A/458/569 Blakeslee J.: J/ApJ/621/53 J/ApJS/150/1 Blakeslee J.P.: J/AJ/108/2087 J/AJ/125/1817 J/AJ/128/1990 J/AJ/131/143 J/ApJ/591/850 J/ApJ/606/L25

J/ApJ/612/202 J/ApJ/613/279 J/ApJ/644/30 J/ApJ/661/750 J/ApJS/153/223 J/ApJS/154/509 J/MNRAS/327/1004 Blanc G.: J/A+A/389/149 Blanc G.A.: J/AJ/134/1103 Blancard C.: J/A+A/439/387 J/A+AS/105/77 Blanchard A.: B/bax J/A+A/342/1 J/A+A/361/429 J/A+AS/130/323 Blanco A.: J/A+A/330/505 Blanco B.M.: J/AJ/114/2596 J/AJ/89/636 Blanco C.: J/A+AS/125/399 Blanco J.: J/A+A/366/868 J/A+A/416/811 J/AJ/127/1181 Blanco V.M.: II/4 J/AJ/114/2596 BLANCO V.M.: J/AJ/80/1011 Blanco V.M.: J/AJ/89/636 J/ApJ/120/129 J/ApJ/125/195 J/ApJ/353/494 Blandford R.: J/AJ/120/2190 J/ApJ/471/L5 J/ApJ/538/29 J/ApJS/127/1 J/ApJS/159/41 Blandford R.D.: VIII/72 Bland-hawthorn J.: J/AJ/130/1065 Bland-Hawthorn J.: J/ApJS/165/307 J/MNRAS/329/227 J/MNRAS/329/87 J/MNRAS/333/100 Bland-hawthorn J.: J/MNRAS/350/1485 J/MNRAS/352/939 Bland-Hawthorn J.: J/MNRAS/372/425 Blanton M.R.: J/AJ/123/567 J/AJ/126/2579 Blaszkiewicz: J/A+A/413/233 Blaufuss E.: J/other/APh/26.282 Blaylock M.: J/ApJ/636/1098 BLEACH R.D.: VII/80 Blecha A.: J/A+A/268/419 J/A+A/419/533 Bleeker J.A.M.: J/A+A/420/135 Bliton M.: J/AJ/118/2014 Blitz L.: J/A+A/275/67 J/ApJ/428/693 J/ApJS/149/343 BLITZ L.: VII/50 VII/76 Bloch J.: J/AJ/119/1901 J/AJ/127/2436 BLOECKER T.: J/A+A/299/755 Bloeckera T.: J/A+A/339/123 Bloemen H.: J/A+AS/143/145 Blom J.J.: J/A+AS/143/145 Blommaert J.: J/A+A/372/173

Blommaert J.A.D.L.: J/A+A/357/225 J/A+A/362/215 J/A+A/406/51 J/A+A/443/143 J/ApJ/552/289

J/MNRAS/376/313 Blomme R.: J/A+A/395/499 Blommeart J.A.D.L.: II/243 Blondel P.F.C.: J/A+AS/124/33 Blondin S.: J/AJ/130/2453 J/AJ/131/1648 J/ApJ/602/571 Bloom J.S.: J/ApJ/608/L93 Bloom S.D.: J/ApJS/123/79 Blum R.D.: J/AJ/132/271 J/ApJ/470/864 J/ApJ/597/323 Blundell K.M.: J/MNRAS/307/293 J/MNRAS/373/1531 J/MNRAS/375/1349 Boboltz D.A.: J/A+A/470/191 Bobylev V.V.: I/295 J/AZh/83/821 J/PAZh/32/906 Boccacci P.: J/AJ/129/1534 Bocchialini K.: J/A+A/313/949 Bochanski J.J.: J/AJ/130/1871 J/AJ/131/1674 Bochkarev N. G.: J/ApJ/510/659 Bochkarev N.G.: J/A+A/371/79 J/A+A/376/775 J/A+A/422/925 J/ApJS/115/185 J/PAZh/27/83

J/PAZh/27/942 Bock H.: J/ApJS/115/185 Bockelee-Morvan D.: J/A+A/393/1053 Bocsa G.: J/other/RoAJ/10.51 Bodaghee A.: J/A+A/411/L59 J/A+A/467/585 J/ApJ/607/L33 J/ApJ/636/765 Bode M.F.: III/174 BODINI L.: J/A+AS/112/347 Boeche C.: J/A+A/415/145 Boehm A.: J/A+A/398/49 J/A+A/418/885 J/A+A/420/97 J/A+A/421/41 J/A+A/422/907 Boehnhardt H.: J/A+A/380/347 Boehringer H.: J/A+A/420/61 J/A+A/425/367 J/A+A/461/397 J/AJ/130/968 J/MNRAS/301/881 Boer M.: J/AJ/133/1470 J/ApJS/120/399 J/other/GCN/1070.1 J/other/IBVS/5767 Boersma C.: J/AJ/130/1977 Boersma D.J.: J/other/APh/26.282 Boerst H.: J/A+A/390/39 Boerst H.-G.: J/A+A/406/L9 J/A+A/410/813 Boeser S.: J/other/APh/26.282 Boeshaar P.C.: J/ApJS/43/379 Boffi F.: J/A+A/465/481 Boffi F.R.: J/A+AS/138/253 Boffin H.M.J.: J/A+A/456/599 J/A+AS/102/361 Bogdanov S.: J/ApJ/625/796 Bogdanovich P.: J/A+A/433/745 J/A+A/455/1157 Bogdantsov A.V.: J/A+AS/139/545 Boggess A.: J/AJ/116/941 Bogun S.: J/A+A/359/865 Bohigas J.: J/MNRAS/367/513 Bohlender D.A.: J/A+A/361/991 J/A+A/458/569 BOHLIN R. C.: III/167

Bohlin R.C.: III/157 J/A+AS/105/305 J/AJ/108/1350 J/AJ/112/601 J/AJ/113/1011 J/AJ/116/180

J/AJ/128/3053 J/AJ/131/866 BOHLIN R.C.: J/ApJ/413/604 J/ApJ/425/122 Bohlin R.C.: J/ApJ/438/181 J/ApJ/446/622 J/ApJS/104/287 J/ApJS/98/595 Bohlsen T.: J/MNRAS/335/1085 Bohm C.: J/other/APh/26.282 Bohn R.K.: J/ApJS/162/428 Bohnenstengel H.D.: VIII/68 Bohringer H.: J/A+A/423/449 J/AJ/127/1811 J/ApJ/514/148 J/ApJ/518/69 J/ApJ/534/114 J/ApJ/617/192

J/ApJ/622/187 J/ApJS/129/435 J/ApJS/140/239 J/ApJS/149/29 J/ApJS/162/304 J/MNRAS/274/75 J/MNRAS/281/799

Boissier S.: J/ApJ/633/871 J/ApJS/164/38 Boisson C.: J/A+AS/100/583 J/A+AS/116/239 J/A+AS/117/93 Bojahr H.: J/A+A/390/39 J/A+A/406/L9 J/A+A/410/813 BOK B.J.: II/119 VI/1 Boker T.: J/AJ/127/105 Boksenberg A.: J/A+A/287/769 J/A+A/288/413 Bolatto A.: J/ApJ/517/282 Boldt E.A.: J/ApJ/253/485 J/ApJS/96/303 BOLDT E.A.: VII/181 Boller T.: IX/32 BOLLER T.: J/A+A/281/355

Boller T.: J/A+A/368/797 J/A+A/378/30 J/A+A/476/1191 J/A+AS/127/145 J/A+AS/129/87

J/AJ/126/2209 BOLLER T.: VII/161 Boller Th.: IX/10A IX/29 Bolt G.: J/A+A/465/331 Bolte M.: J/A+A/390/961 BOLTE M.: J/AJ/107/1764 J/AJ/107/679

Bolte M.: J/AJ/112/2408 J/AJ/117/247 J/AJ/121/148 J/AJ/128/2402 J/AJ/128/3019 J/AJ/97/1360

J/ApJ/518/262 J/ApJ/527/199 J/ApJ/608/243 J/ApJ/611/323 J/ApJ/628/729 J/ApJS/120/265 J/PASP/109/883 Bolton C.T.: J/A+A/337/183 J/A+A/458/569 Bolton R.C.: J/MNRAS/354/485 J/MNRAS/379/1442 Boltwood P.: J/A+AS/129/577 Bolz O.: J/A+A/390/39 J/A+A/406/L9 J/A+A/410/813 Bolzicco G.J.: III/204 Bolzonella M.: J/A+A/403/857 J/A+A/428/1043 J/A+A/457/79 Bomans D.J.: J/A+A/328/167 J/A+A/441/905 J/A+A/458/341 BOMANS D.J.: J/A+AS/112/367 Bomans D.J.: J/A+AS/123/455 Bonanno A.: J/A+A/459/557 Bonanno G.: J/A+A/377/123 J/A+A/420/683 J/AJ/121/327 Bonanos A.Z.: J/AJ/126/175 J/AJ/127/861 J/ApJ/591/L111 J/ApJ/626/L5 Bonatto C.: III/219 J/A+A/450/105 BONAZZOLA S.: J/A+A/291/155 J/A+AS/108/455 Bond H.E.: J/A+A/434/1107 J/AJ/117/247 J/AJ/121/318 J/ApJS/120/265 Bond I.: J/MNRAS/331/609 Bond I.A.: J/MNRAS/330/137 Bond J.C.: J/MNRAS/370/163 Bondar A.: J/ApJ/634/451 Bondar N.I.: J/AZh/79/542 J/PAZh/26/525 Bondarenko I.I.: V/120 Bondi M.: J/A+A/379/755 J/A+A/403/857 J/A+A/428/1043 J/A+A/457/79 J/A+A/463/519 J/A+AS/120/89

J/AJ/128/1974 Bondin S.: J/ApJ/613/200 Bonfanti C.: J/A+A/417/499 Bonfanti: J/A+AS/139/483 BONFANTI P.: J/A+A/297/28 Bonfils X.: J/A+A/471/L51 J/A+A/474/293 J/A+A/475/1125 Bongiorno A.: J/A+A/428/1043 J/A+A/457/79 Bongiorno B.J.: J/AJ/133/2258 Bongiovanni A.: J/AJ/127/1158 J/MNRAS/359/930

Bonifacio P.: J/A+A/319/593 J/A+A/352/117 J/A+A/403/1105 J/A+A/414/503 J/A+A/416/1117

J/A+A/464/201 J/A+A/465/815 Bonito N.: V/107 Bonneau D.: J/A+A/312/879 J/A+A/378/954 Bonnefond P.: J/A+AS/118/277 BONNET R.: IV/11 IV/8 Bonnoli G.: J/A+A/468/807 Bono G.: J/A+AS/121/327 J/AJ/128/1167 J/AJ/129/2714

Bontemps S.: J/A+A/350/883 J/A+A/357/219 J/A+A/359/113 J/A+A/372/173 J/A+A/372/L41

J/A+A/420/945 J/A+A/421/623 J/A+A/458/789 J/A+A/476/1243 Boogert A.C.A.: J/ApJ/606/L73 Bookbinder J.: J/ApJ/315/687 J/ApJ/348/253 Boonyarak C.: J/ApJ/616/562 Booth J.: J/ApJ/344/704 Booth R.: J/A+AS/126/3 Booth R.S.: J/A+A/276/25 J/A+A/376/1123 J/A+A/403/1095 J/A+AS/115/439 J/A+AS/132/211 BOOTH R.S.: J/A+AS/93/121 BOPP B.W.: J/A+AS/103/183 Bopp B.W.: J/A+AS/139/513 Borch A.: II/253 Borczyk W.: J/A+A/474/473 Borde P.: J/A+A/393/183 J/A+A/433/1155 BORDE S.: III/165 BORDEAUX: I/66 Border J.: J/PASJ/52/997 Borgani S.: J/A+A/432/381 J/A+A/451/457 J/ApJS/139/369 Borghesi A.: J/A+A/330/505 Borgonovo L.: J/A+A/452/975 Borguet B.: J/A+A/478/321 Borisov G.V.: III/209 Borisov N.: J/A+A/422/925 J/A+A/476/1191 Borisov N.V.: J/AZh/79/309 J/PAZh/28/764

Borissova J.: J/A+A/400/917 J/A+A/423/97 J/A+A/474/121 J/A+AS/110/1 J/A+AS/121/499

J/AJ/119/2274 J/AJ/124/3157 J/ApJ/640/L43 Borkova T.V.: J/AZh/79/510 J/AZh/82/453 J/PAZh/30/173 Borkovits T.: J/A+A/349/515 J/A+A/403/637 Borkowski J.: J/A+A/469/799 Borkowski V.: J/A+A/474/473 Borne K.D.: J/AJ/130/2043 Bornmann P.: J/AJ/118/2014 Boroski W.N.: J/AJ/121/2331 J/AJ/123/567 J/AJ/126/2579 Boroson T.: J/AJ/130/381 Boroson T.A.: J/AJ/121/66 J/AJ/123/1292 J/AJ/127/1943 J/AJ/130/2571 J/ApJ/585/647 Boroumand S.: J/A+A/465/331 Borovicka J.: J/A+AS/103/83 Borra E.F.: J/ApJ/509/309 BORRA. E.F.: J/MNRAS/247/182

Borsenberger J.: J/A+A/349/L69 J/A+A/362/215 J/A+A/363/62 J/A+A/375/275 J/A+A/380/590

J/A+A/387/1 J/A+A/401/959 J/A+A/423/381 J/A+AS/135/133 VII/242 Borun B.C.P.: J/other/SerAJ/166.13Short Borys C.: J/MNRAS/358/149 J/MNRAS/372/1621 J/MNRAS/376/1073 Bos M.: J/A+A/432/657 Bosch G.: J/A+A/341/98 Bosch G.L.: J/MNRAS/355/1237 Boschi F.: J/A+A/406/995 J/A+A/418/869

Boschin W.: J/A+A/396/397 J/A+A/397/431 J/A+A/416/839 J/A+A/444/685 J/A+A/449/461

J/A+A/455/45 J/A+A/467/37 J/A+A/469/861 Bosch-Ramon V.: J/A+A/472/557 BOSELLI A.: J/A+A/283/12 Boselli A.: J/A+A/327/522 J/A+A/343/86 J/A+A/343/93 J/A+A/361/863 J/A+A/384/383 J/A+A/396/449

J/A+A/397/421 J/A+A/401/471 J/A+A/402/37 J/A+A/406/867 J/A+A/411/381 J/A+A/417/499 J/A+A/419/47 J/A+A/425/429 J/A+A/430/411 J/A+A/446/839 J/A+A/449/929 J/A+AS/104/271 J/A+AS/112/257 J/A+AS/116/193 J/A+AS/120/489 J/A+AS/120/521 J/A+AS/121/507 J/A+AS/125/293 J/A+AS/142/65 J/A+AS/142/73 J/ApJ/633/871 J/ApJS/164/38 J/MNRAS/332/422

Bosh A.S.: J/AJ/103/983 Bosio: J/A+A/327/61 Bosio S.: J/A+AS/121/119 Bosler T.L.: J/MNRAS/378/318 Bosma A.: J/A+A/385/816 J/A+A/421/595 Bosq F.: I/272 Boss B.: I/113A BOSS B.: I/25 I/8 Bossi M.: J/A+A/363/29 J/A+A/367/759 Bot C.: J/A+A/471/103 Botashev A.M.: J/AZh/76/83 J/other/BSAO/46.28 Bothmer V.: J/A+A/422/337 Bothun G.: J/A+A/428/823 BOTHUN G.D.: J/AJ/106/530 Bothun G.D.: J/ApJ/613/914 J/ApJS/105/209 J/ApJS/128/99 Botner O.: J/other/APh/26.282 Botschwina P.: J/A+A/425/767 Bottemiller R.L.: II/83 Botti L.C.L.: J/A+AS/116/157 Bottinelli L.: J/A+A/113/61 J/A+A/430/373 J/A+AS/102/57 J/A+AS/130/333 J/A+AS/91/371 VII/119

VII/136 VII/238 Bottinelli S.: J/MNRAS/355/97 Bottini D.: J/A+A/403/857 J/A+A/428/1043 J/A+A/457/79 Bouche N.: J/MNRAS/354/L25 J/MNRAS/371/495 Bouchet F.R.: J/A+A/372/364 Bouchet P.: J/A+A/365/545 J/A+AS/119/547 J/A+AS/125/399 J/AJ/104/28 J/ApJ/615/887 Bouchez A.H.: J/AJ/112/1197 Bouchta A.: J/other/APh/26.282

Bouchy F.: J/A+A/424/L31 J/A+A/466/743 J/A+A/474/293 J/A+A/475/1125 J/A+A/476/1347

J/A+A/478/197 BOUDIN F.: VI/63 Boughaleb H.: J/A+AS/124/143

Boulanger F.: J/A+A/276/25 J/A+A/350/883 J/A+A/357/219 J/A+A/372/173 J/A+A/420/945

J/A+A/421/623 J/A+A/471/103 Boulesteix J.: J/A+A/330/881 BOULESTEIX J.: J/A+AS/113/35 Boulesteix J.: J/A+AS/137/419 J/MNRAS/328/1181 Boumis P.: J/A+A/402/1157 J/A+A/456/599 BOUNATIRO L.: VI/60 Bouquet A.: J/A+A/389/149 J/A+A/421/509 Bourban G.: J/A+A/406/L23 J/A+A/411/L59 Bourgeois J.: J/A+AS/125/399 Bourget P.: J/A+A/425/1107 J/A+A/454/683 Bourgois G.: J/A+A/393/1053 Bourke T.: J/A+A/413/1177 Bourke T.L.: J/A+A/475/549 Bournaud F.: J/A+A/438/507 Boutelier T.: J/A+A/463/1017 BOUVIER J.: J/A+A/299/89

Bouvier J.: J/A+A/323/139 J/A+A/335/183 J/A+A/349/619 J/A+A/375/989 J/A+A/395/813 J/A+A/400/891 J/A+A/409/169 J/A+A/436/853 J/A+A/448/189 J/A+A/450/147 J/A+A/458/461 J/A+A/461/183 J/A+A/463/1017 J/A+A/468/353 J/A+A/468/379 J/A+A/471/499

BOUVIER J.: J/A+AS/101/485

Bouvier J.: J/A+AS/120/127 J/ApJ/546/1006 J/ApJ/593/1093 J/ApJ/610/1064 J/MNRAS/370/954

J/MNRAS/377/741 Bouwens R.: J/AJ/130/1 J/ApJ/621/53 J/ApJS/150/1 Bouwens R.J.: J/AJ/128/1990 J/ApJ/606/L25 J/ApJ/612/202 J/ApJ/624/L5 J/ApJ/644/30 J/ApJS/157/175 Bouwman J.: J/PASP/118/1690 Bouy H.: J/AJ/126/1526 Bowell E.: B/astorb J/AJ/109/1363 J/AJ/119/2008 Bower G.C.: J/AJ/127/3399 Bower R.: J/MNRAS/311/456 Bower R.G.: J/ApJ/549/820 J/MNRAS/335/10 J/MNRAS/337/256 J/MNRAS/358/233 Bowers C.W.: J/ApJ/532/308 Bowers T.P.: J/ApJ/565/511 Bowyer S.: II/203A II/203A IX/35 IX/35 J/A+A/250/370 J/A+A/321/531 J/A+AS/127/1 J/AJ/107/751

J/ApJS/111/143 J/ApJS/93/569 J/ApJS/96/461 J/MNRAS/313/641 J/MNRAS/324/580 J/MNRAS/332/441 J/PASP/109/998

Boyarchuk A.A.: J/AZh/72/864 J/AZh/73/862 J/AZh/75/586 J/AZh/77/96 J/AZh/78/807 J/AZh/79/909

J/AZh/80/704 J/AZh/81/658 J/AZh/82/601 Boyarchuk M.E.: J/AZh/72/864 J/AZh/73/862 J/AZh/75/586 J/AZh/77/96 J/AZh/79/909 Boyce E.: J/AJ/130/896 Boyce P.J.: J/AJ/123/873 J/AJ/124/1954 J/AJ/124/690 J/AJ/128/16 J/MNRAS/337/641 J/MNRAS/342/738

J/MNRAS/346/787 J/MNRAS/355/1303 Boyce P.W.: J/A+AS/114/253 Boyd D.: J/other/IBVS/5526 Boyd D.M.: J/ApJS/99/551 Boyle B. J.: J/ApJS/122/355 Boyle B.: J/AJ/131/2383

Boyle B.J.: II/268 J/AJ/121/2381 J/AJ/130/1358 J/ApJ/505/L7 J/MNRAS/243/1 J/MNRAS/251/482 J/MNRAS/268/305 J/MNRAS/276/33 J/MNRAS/306/592 J/MNRAS/328/805 J/MNRAS/335/673 J/MNRAS/349/135 VII/223 VII/241

BOYLE R.P.: II/130 Boyle R.P.: J/A+A/388/149 J/A+A/458/767 J/BaltA/11/465 J/BaltA/12/323 J/BaltA/13/1 J/BaltA/13/555

J/BaltA/15/483Short J/BaltA/7/369 Bozic H.: J/A+A/304/235 J/A+A/312/879 J/A+A/324/965 J/A+A/328/551 J/A+A/350/566 J/A+A/367/236

J/A+A/387/580 J/A+A/405/1087 J/A+A/408/611 J/A+A/415/289 J/A+A/416/669 J/A+A/419/607 J/A+A/455/1037 J/A+A/463/1061 J/A+A/463/233 J/A+A/464/263

BOZIC H.: J/A+AS/107/403 Bozic H.: J/A+AS/125/75 Bozyan E.P.: J/ApJS/82/1 Bozyan F.A.: VIII/42 B.P.Artamonov: J/A+A/420/447 Braatz A.: J/A+A/423/983 Bradley T.R.: J/A+A/459/L13 VI/112 BRAEUER H.J.: VI/58 Braeuer H.-J.: VI/68 Braeuninger H.: IX/10A

Bragaglia A.: J/A+A/354/169 J/A+A/364/102 J/A+A/413/343 J/A+A/430/603 J/A+A/440/901 J/A+A/455/271 J/A+A/464/927 J/A+A/464/939 J/A+A/464/953 J/A+A/464/967 J/A+A/465/185 J/A+AS/102/397 J/AJ/121/327 J/MNRAS/299/834 J/MNRAS/309/739 J/MNRAS/363/734 J/MNRAS/375/1381

Bragaglia: J/A+A/421/937 Bragg A.E.: J/AJ/131/527 Braine J.: J/A+A/411/381 J/A+AS/114/21 Brainerd J.J.: IX/20A Braito V.: J/A+A/428/383 Braje: J/A+A/386/1001 Bramich D.M.: J/MNRAS/362/1006 J/MNRAS/377/741 VI/112 BRANCEWICZ H.K.: II/150A Branch D.: J/AJ/131/1648 J/ApJ/573/306 Branchesi M.: J/A+A/446/97 J/A+A/462/449 Brand J.: J/A+A/275/67 J/A+A/308/573 J/A+A/336/339 J/A+A/368/845 J/A+A/430/549 J/A+A/432/921

J/A+A/447/221 J/A+A/464/909 J/A+AS/103/541 J/A+AS/80/149 J/A+AS/98/589 Brand K.: J/ApJ/641/140 J/ApJS/161/9 J/MNRAS/344/283 J/MNRAS/357/1231 Brandi E.: J/A+A/343/841 BRANDIMARTI A.: J/A+AS/104/379 Brandl B.: J/ApJ/466/254 Brandl B.R.: J/ApJ/635/280 Brandner W.: J/AJ/126/1526 J/ApJS/155/123 J/ApJS/166/549 Brandt J.C.: J/AJ/116/941 Brandt S.: J/A+A/427/87 J/A+AS/129/1 Brandt W.N.: J/AJ/122/2810 J/AJ/124/2351 J/AJ/126/2579 J/AJ/126/539 J/AJ/126/632 J/AJ/130/387

J/ApJ/554/742 J/ApJ/589/225 J/ApJ/613/1179 J/ApJ/640/167 J/ApJS/161/21 J/MNRAS/311/456 Branduardi-Raymont G.: J/MNRAS/274/1194 J/MNRAS/311/456 Brassard P.: J/A+A/459/557 Bratschi P.: J/A+A/318/416 J/A+A/320/74 Bratsolis E.: J/A+A/423/919 Braude et al.: VIII/41 VIII/80 Brault J.W.: J/ApJS/94/221 Braun J.: J/other/APh/26.282 Braun J.M.: II/232 J/A+A/328/167 J/A+AS/145/223 J/ApJS/115/185 Braun M.: J/A+AS/117/39 VII/196 Braun R.: J/A+A/391/159 J/A+A/406/829 J/A+A/417/421 J/A+AS/92/625 J/AJ/123/873 J/ApJ/623/181 Brauninger H.: IX/29 Bravo A.: J/A+AS/123/1 J/A+AS/123/59 Brayton A.M.: J/ApJ/81/187 Breger M.: III/48B Bregman J.D.: J/AJ/110/275 J/AJ/112/2274 J/AJ/112/241 Bregman J.N.: J/ApJ/405/94 J/ApJS/157/59 BREHIN S.: J/A+AS/109/447 Breinhorst R.A.: J/A+A/452/959 Brekke P.: J/A+A/375/591

Bremer M.: J/A+A/444/365 J/A+A/456/791 J/A+A/476/1191 J/A+A/476/L17 J/A+AS/134/401

J/ApJ/591/640 J/ApJS/107/175 J/MNRAS/339/913 J/MNRAS/379/1343 J/MNRAS/382/279 Bremer M.A.R.: J/A+AS/124/259 Bremer M.N.: J/A+A/339/34 J/A+AS/124/259 Brenneman L.: J/MNRAS/343/665 Bresolin F.: J/A+A/440/783 J/A+A/441/981 J/AJ/128/1167 J/AJ/128/2815 J/AJ/130/652 J/ApJ/507/655

J/ApJ/510/104 J/ApJ/512/48 J/ApJ/521/155 J/ApJ/523/540 Bressan A.: J/A+A/311/361 J/A+A/335/823 J/A+A/345/419 J/A+A/433/497 J/A+A/434/149

J/A+A/445/79 J/A+A/463/455 J/A+AS/106/275 BRESSAN A.: J/A+AS/112/35 Bressan A.: J/A+AS/130/251 J/A+AS/130/513 J/A+AS/141/371 J/MNRAS/364/146 VI/96 Bretagnon P.: J/A+A/329/329 J/A+A/408/387 BRETAGNON P.: VI/66 VI/81 VI/81 Bretherton D.E.: J/MNRAS/379/663 BRETT J.M.: J/A+AS/109/263 Brewer J.: J/AJ/127/2771 Brewer J.P.: J/AJ/109/2480 J/AJ/112/491 J/AJ/129/729 Brewer M.-M: J/AJ/131/431 Brewington H.J.: J/A+A/444/643 J/AJ/131/1674 J/AJ/132/676 BREYSACHER J.: J/A+AS/76/101 Brian D. Mason: I/281 I/301

Briceno C.: J/AJ/127/1158 J/AJ/128/805 J/AJ/129/856 J/AJ/129/907 J/AJ/130/188 J/AJ/131/527

J/ApJS/134/103 Bridges R.J.: J/MNRAS/300/303 Bridges T.: J/AJ/124/404 J/MNRAS/329/227 J/MNRAS/329/87 J/MNRAS/333/100 J/MNRAS/350/1485

J/MNRAS/352/939 J/MNRAS/355/608 J/MNRAS/372/425 Bridges T.J.: J/A+A/372/391 J/AJ/123/2490 J/ApJ/559/812 J/ApJS/137/279 J/MNRAS/312/442

J/MNRAS/369/120 J/MNRAS/369/97 Bridle A.H.: J/AJ/77/405 J/AJ/98/419 Bridzius A.: J/PASJ/56/1025 Briel U.: IX/10A IX/29 J/A+A/316/147 Briel U.G.: J/A+A/278/379 J/A+A/419/47 J/A+AS/126/509 J/ApJ/617/192 J/ApJS/149/29 J/ApJS/162/304

J/MNRAS/281/799 Brieva E.: J/PASP/114/233 Briggs J.W.: J/AJ/121/2308 J/AJ/123/567 J/ApJS/163/80 Briggs K.R.: J/A+A/468/353 J/A+A/468/379 J/MNRAS/326/1161 J/MNRAS/367/781 Briggs M. S.: J/ApJS/120/399 Briggs M.S.: IX/20A J/ApJS/126/19 Bright J.: J/AJ/127/2674 Briley M.M.: J/AJ/127/1579 J/AJ/127/1588 J/AJ/130/1177 Brinch: J/A+A/475/915 Brinchmann J.: J/A+A/448/893 J/A+A/462/535 J/ApJ/499/112 Brinkman J.: J/A+A/444/643 J/ApJS/145/245 J/ApJS/165/1 Brinkmann J.: J/A+A/423/449 J/AJ/121/2308 J/AJ/121/2331 J/AJ/122/503 J/AJ/123/1807 J/AJ/123/3409

J/AJ/123/567 J/AJ/124/1651 J/AJ/125/1711 J/AJ/125/1817 J/AJ/126/1023 J/AJ/126/1720 J/AJ/126/2125 J/AJ/126/2209 J/AJ/126/2579 J/AJ/127/2838 J/AJ/127/3034 J/AJ/127/3553 J/AJ/127/899 J/AJ/128/426 J/AJ/129/2542 J/AJ/130/873 J/AJ/131/1 J/AJ/131/1288 J/AJ/131/1674 J/AJ/131/2018 J/AJ/131/2722 J/AJ/131/49 J/AJ/131/84 J/AJ/132/676 J/ApJ/588/824 J/ApJ/598/260 J/ApJ/607/426 J/ApJ/633/844 J/ApJS/148/243 J/ApJS/155/257 J/ApJS/163/80 J/PASJ/55/771

Brinkmann J.V.: J/ApJ/633/638 J/ApJ/641/78 BRINKMANN W.: J/A+A/281/355 J/A+A/285/812

Brinkmann W.: J/A+A/309/81 J/A+A/319/413 J/A+A/323/739 J/A+A/330/108 J/A+A/356/445 J/A+A/363/141 J/A+AS/106/303 J/A+AS/109/147 J/A+AS/122/235 J/A+AS/134/221 J/ApJ/450/51 J/ApJ/525/127

BRINKMANN W.: J/ApJS/93/145 Brinks E.: J/A+AS/61/451 J/AJ/121/182 Briot D.: J/A+AS/137/451 Briquet M.: J/A+A/457/651 J/A+A/463/243 Brisbin D.: J/MNRAS/367/1609 Brisken W.F.: J/AJ/121/1192 Britzen: J/A+A/364/391 Britzen S.: J/A+A/472/763 Broadbent A.: J/MNRAS/300/417 Broadhurst T. J.: J/ApJS/122/355 Broadhurst T.: J/A+A/392/395 J/A+A/437/883 J/ApJ/499/112 J/ApJ/621/53 J/MNRAS/273/157

Broadhurst T.J.: J/A+A/432/381 J/AJ/128/1990 J/ApJ/606/L25 J/ApJ/612/202 J/ApJS/150/1

J/MNRAS/235/827 Brocato E.: J/A+A/351/526 BROCATO E.: J/AJ/102/137 Brocato E.: J/ApJS/71/25 Brocksopp C.: J/ApJ/636/1002 Broderick J.J.: J/AJ/109/2318 J/ApJS/129/547 BRODERICK J.J.: VII/97 VIII/26 Broderick J.J.: VIII/6 VIII/65 Brodie J.P.: J/A+A/443/413 J/AJ/119/727 J/AJ/123/2490 J/AJ/131/814 J/ApJS/102/29 J/MNRAS/327/1116

J/MNRAS/366/1230 Brodwin M.: J/ApJ/639/816 J/MNRAS/350/785 Broeders E.: J/A+A/463/243 Broeg C.: J/A+A/468/1039 Brogan C.L.: J/AJ/128/1646 Brogt E.: J/A+A/392/1129 J/A+A/404/L29 Broguiere D.: J/A+A/476/1243 Bromage G.E.: J/A+A/476/1191 J/MNRAS/274/1194 Bronfman L.: J/A+A/327/325 J/A+A/426/97 J/A+A/463/1029 J/A+AS/115/81 J/ApJ/602/850 Bronnikova N.M.: I/290 I/295 J/A+A/396/125 J/A+A/451/901 Brook A.: J/MNRAS/313/641 Brooke T.: J/PASP/118/1690 BROOKE T.Y.: J/AJ/106/656 Brookes M.H.: J/MNRAS/346/627 J/MNRAS/366/1265 Brooks D.H.: J/A+A/432/1063 Broos P.: J/AJ/120/1426 J/ApJ/574/258 J/ApJS/160/353 J/ApJS/163/306 Broos P.S.: J/AJ/122/2810 J/AJ/126/539 J/AJ/131/2164 J/ApJ/554/742 J/ApJ/589/225 J/ApJ/613/1179

J/ApJS/160/319 Brosch N.: J/A+A/366/771 J/A+A/374/907 J/A+A/419/469 J/A+A/442/109 J/A+AS/142/417 J/AJ/117/206 J/AJ/130/2613 J/ApJS/111/143 J/MNRAS/313/641 J/MNRAS/324/580 J/MNRAS/332/441 J/MNRAS/357/645

Brosche P.: J/A+A/312/74 BROSCHE P.: J/A+AS/104/161 Brosius J.W.: J/A+A/370/1071 J/ApJ/543/1016 Brotherton M.S.: J/A+A/356/445 J/ApJ/505/L7 J/ApJS/126/133 J/ApJS/135/227 J/MNRAS/375/989 Brott I.: J/A+A/442/281 BROTZMAN L.E.: V/51 VII/63 Brough S.: J/MNRAS/369/1351 J/MNRAS/370/1223 Brouty M.: VII/237 Brown A.: J/A+A/310/173 BROWN A.G.A.: J/A+A/289/101

Brown A.G.A.: J/A+A/331/81 J/A+A/428/149 J/A+A/430/137 J/A+A/464/581 J/AJ/117/354

J/AJ/124/404 Brown B.H.: J/A+AS/106/333 BROWN D.R.: VIII/45 Brown E.F.: J/ApJ/578/405 Brown J.A.: J/AJ/118/1245 Brown J.C.: J/A+A/477/593 J/ApJS/145/213 Brown J.M.: J/A+A/406/L47 Brown L.W.: II/71 Brown M.A.: J/AJ/129/413 J/AJ/130/1658 Brown M.E.: J/AJ/112/1197 J/ApJ/564/421 Brown M.J.I.: J/AJ/121/2381 J/AJ/127/213 J/ApJ/639/816 J/ApJ/641/140 J/ApJS/161/9 Brown P.J.F.: J/A+A/315/95 Brown R.A.: J/AJ/128/1990 J/AJ/129/2294 J/ApJ/612/202 J/ApJ/621/53 J/ApJS/150/1 Brown R.D.: J/A+A/459/317 Brown T.M.: J/A+A/410/1051 J/A+A/476/1347 J/AJ/119/486 J/AJ/127/2738 J/AJ/130/1693 J/ApJ/504/113

J/ApJ/532/308 J/ApJ/559/1060 J/ApJ/562/368 J/ApJ/582/1123 J/ApJ/634/625 Brown W.R.: J/AJ/122/2893 J/AJ/122/714 J/AJ/126/1362 J/AJ/130/1097 J/AJ/131/527

Browne I.W.A.: J/A+A/472/763 J/A+A/474/1093 J/ApJS/100/1 J/ApJS/95/345 J/MNRAS/326/1455

J/MNRAS/329/700 J/MNRAS/376/371 J/MNRAS/377/430 VIII/60 VIII/72 Browne S.E.: J/AJ/130/825 BROWNING J.C.: J/AJ/105/2265 Browning L.: J/PASP/109/289 Brownstein J.R.: J/ApJ/636/721 Bruch A.: J/A+A/332/586 J/A+A/374/1056 J/A+AS/70/481 J/A+AS/93/463 Bruenken S.: J/A+A/459/317 Bruevich V.V.: J/A+A/420/447 Bruggen M.: J/MNRAS/379/260 Brugger H.R.: J/A+AS/121/247 J/MNRAS/306/279 Brumberg V.A.: J/A+A/408/387 Brunner H.: IX/32 J/A+A/311/384 J/A+A/428/383 J/A+A/469/27 J/A+A/476/1191 J/A+AS/114/465 Brunner R.: J/AJ/121/2308 J/AJ/122/503 J/AJ/123/567 J/AJ/126/2579 J/ApJS/159/41

Brunner R.J.: J/AJ/121/31 J/AJ/125/2064 J/AJ/128/1017 J/AJ/131/1 J/AJ/131/1923 J/ApJ/633/638

J/ApJ/641/78 J/ApJS/155/257 J/ApJS/163/80 J/MNRAS/372/425 Bruno P.: J/AJ/121/327 Brunswig W.: J/A+AS/102/611 Brunt C.M.: J/A+A/399/1083 J/ApJ/626/195 J/ApJS/144/47 Bruntt H.: J/A+A/456/651 J/A+A/467/1215 J/A+A/471/515 J/A+AS/140/135 J/PASP/115/413 Brunzendorf J.: J/A+A/349/448 J/A+A/357/471 J/A+A/369/511 J/A+AS/137/305 J/A+AS/139/141 Brusa M.: J/A+A/409/79 J/A+A/432/69 J/A+A/457/501 J/A+A/466/31 Brusasco M.A.: J/other/RMxAA/39.89 BRUZUAL A.G.: J/A+A/297/61 Bruzual A.G.: J/ApJ/611/270 Bruzual G.: III/232 J/AJ/127/1158 J/MNRAS/359/930 Bryden G.: J/ApJ/620/1010 J/ApJ/636/1098 BRYJA C.: J/PASP/104/553 B.T.: J/A+A/386/1001 Buat V.: J/A+A/313/129 J/A+A/327/522 J/ApJS/164/38 J/MNRAS/380/986 Bubar E.J.: J/AJ/132/161 Bucciarelli B.: I/176 J/A+A/448/579 J/A+AS/121/119 Bucciarelli et al.: II/272 Bucharest: I/31 Buchhorn M.: J/ApJS/81/413 Buckle J.: J/MNRAS/373/781 Buckley D.: J/ApJ/459/606 Buckley D.A.H.: J/MNRAS/274/1194 Buckley J.H.: J/AJ/131/527 Budavari T.: J/AJ/127/131 BUDDING E.: J/A+A/291/110 Budding E.: J/A+A/417/263 V/46 Buell J.F.: J/A+A/432/861 Buemi C.S.: J/A+A/327/1114 Buhler S.: J/A+A/465/375 Buil C.: J/A+A/387/580 J/A+A/408/305 J/A+A/416/669 Bujarrabal V.: J/A+A/285/247 J/A+A/319/235 J/A+A/329/219 Bukhmastova Yu.L.: I/292 J/AZh/78/675 J/PAZh/29/253 J/PAZh/33/403 Bukvic S.: J/A+A/477/967 J/A+AS/102/607 Bulder H.J.J.: J/A+AS/125/399 Bulut I.: J/A+A/387/240 J/A+A/417/263 J/MNRAS/378/179 Bunclark P.S: J/MNRAS/259/563 Bunker A.J.: J/AJ/122/598 J/MNRAS/361/525 BUNTON J.D.: VIII/45 BUONANNO R.: J/A+A/290/69

Buonanno R.: J/A+A/310/715 J/A+A/320/757 J/A+A/324/915 J/A+A/330/453 J/A+A/380/478

J/A+A/425/509 J/A+A/465/815 J/A+AS/61/473 BUONANNO R.: J/A+AS/67/327 J/A+AS/84/59 J/AJ/100/1811 J/AJ/102/137 Buonanno R.: J/AJ/109/650 J/AJ/109/663 J/AJ/118/1671 J/AJ/121/916 J/ApJS/71/25 BUONANNO R.: J/MNRAS/252/357 Buonanno R.: J/MNRAS/264/273 BUONANNO R.: J/MNRAS/272/391 Buondi V.: II/227 Buontempo M.E.: I/144 Buote D.A.: J/ApJ/612/848 BURBIDGE G. Astrophys. J. Suppl. Ser. 77: J/ApJS/77/203 BURBIDGE G.: VII/105 Burbidge G.: VII/117 VII/158 VII/178 BURBIDGE G.: VII/27 VII/37 BURBIDGE G.R.: VII/17 VII/33 Burbidge G.R.: VIII/3 BURCHI R.: J/A+AS/101/195 Burchi R.: J/A+AS/125/399 Burdick S.: V/107 Burenin R.: J/ApJ/645/955 Burenkov A.: J/ApJS/115/185 Burenkov A. N.: J/ApJ/510/659 Burenkov A.N.: J/A+A/371/79 J/A+A/376/775 J/A+A/422/925 J/PAZh/27/250 Burford J.: J/AJ/122/2587 Burg R.: IX/22 J/A+A/329/482 J/ApJS/129/435

Burgarella D.: J/A+A/287/769 J/A+A/313/129 J/A+A/324/915 J/A+A/465/1099 J/ApJ/619/L95

J/ApJS/164/38 J/MNRAS/380/986 Burgasser A.J.: J/AJ/125/3302 J/AJ/126/2487 J/AJ/127/2856 J/AJ/128/463 J/AJ/132/891 J/AJ/133/439

J/ApJ/564/421 J/ApJ/637/1067 Burgay M.: J/MNRAS/352/1439 J/MNRAS/372/777 Burgdorf M.: J/A+A/350/883 J/A+A/357/219 J/A+A/372/173 J/A+A/420/945 J/A+A/421/623 Burgess A.M.: J/AJ/131/100 J/AJ/131/114 VIII/16 Burgess C.: J/other/APh/26.282 Burgess P.: J/PASJ/52/997 Burgess T.: J/other/APh/26.282 Burgos-Martin J.: J/MNRAS/311/456 Burigana C.: J/A+A/397/213 J/A+A/410/847 Burke B. F.: J/ApJ/510/64 Burke B.: J/ApJS/75/801 J/AZh/73/835 VIII/52B Burke B.F.: J/AJ/130/2866 J/ApJ/636/1020 J/ApJS/51/67 J/PASJ/52/997 BURKE B.F.: VIII/27 VIII/27 Burke B.F.: VIII/38 VIII/38 VIII/38 VIII/38 VIII/52B VIII/52B Burke C.J.: J/AJ/130/1916 Burke D.J.: J/ApJS/126/209 J/MNRAS/341/1093 Burke M.: J/AJ/110/1186 Burke P.G.: J/A+A/446/1185 Burke V.M.: J/A+A/446/1185 J/A+AS/103/273 Burkert W.: IX/10A IX/29 Burkey J.: VI/76 Burkholder V.: J/ApJ/490/328 Burkhonov O.A.: J/A+A/420/447 BURKI G.: J/A+A/290/609 Burki G.: J/A+A/337/779 J/A+A/341/469 J/A+A/343/847 J/A+A/385/488 J/A+A/395/97 J/A+A/398/1073

J/A+A/401/271 J/A+A/406/373 BURKI G.: J/A+AS/108/25 J/A+AS/108/9 J/A+AS/113/483 Burki G.: J/A+AS/132/367 Burki: J/A+A/454/855 Burleigh M.R.: J/ApJS/147/145

Burles S.: J/AJ/121/2308 J/AJ/121/2331 J/AJ/123/567 J/AJ/125/1817 J/AJ/126/1023 J/AJ/126/2579

J/AJ/131/1 J/ApJS/163/80 Burles S.M.: J/ApJ/639/766 Burley G.S.: J/AJ/132/85 Burley J.: V/31A BURLEY J.: V/6 Burmeister M.: J/A+A/415/273 BURNAGE R.: J/A+AS/110/177 Burnage R.: J/A+AS/124/255 J/A+AS/135/503 J/A+AS/137/273 J/A+AS/137/451 Burnashev V.I.: III/126

Burnet M.: J/A+A/337/779 J/A+A/356/590 J/A+A/375/205 J/A+A/379/999 J/A+A/385/488

J/A+A/388/632 J/A+A/390/267 J/A+A/392/215 J/A+A/415/391 J/A+A/458/895 BURNET M.: J/A+AS/108/9 J/A+AS/113/483 Burnet M.: J/A+AS/127/201 BURNHAM R. Jr.: I/112 Burnham R.Jr: I/79 Burningham B.: J/MNRAS/346/1143 J/MNRAS/356/1583 J/MNRAS/358/341 Burnley A.W.: J/MNRAS/353/601 Burns J.O.: J/AJ/118/2014 J/AJ/120/2269 J/AJ/126/119 J/AJ/126/2740 J/AJ/98/64 J/ApJ/416/36

J/ApJS/80/501 Burns M.S.: J/A+A/437/789 Buromsky M.: J/A+A/438/377 Burrows A.: J/ApJ/564/421 Burrows C.: J/AJ/113/1001 Burrows C.J.: J/AJ/117/1700 J/AJ/128/1990 J/ApJ/612/202 J/ApJ/621/53 J/ApJS/150/1 Burrows D.N.: J/A+A/448/L9 J/AJ/122/2810 J/ApJ/554/742 BURROWS D.N.: VII/43 Burrows: J/PASP/108/682 Bursov N.A.: J/A+A/370/78 Bursov N.N.: J/A+AS/101/447 J/A+AS/87/1 Bursov N.N: J/AZh/74/42 Bursov N.N.: J/AZh/78/922 J/AZh/84/227 J/AZh/84/387 Bursov N.N: J/other/BSAO/40.128 Bursov N.N.: VIII/33 Burstein A.: J/AJ/87/1165 Burstein D.: J/A+A/449/143 J/AJ/112/628 J/AJ/114/1365 J/AJ/114/699 J/AJ/117/2757 J/AJ/126/1286

J/AJ/127/2579 J/AJ/131/866 BURSTEIN D.: J/ApJ/368/54 Burstein D.: J/ApJS/106/1 J/ApJS/109/333 J/ApJS/111/163 J/ApJS/116/1 J/ApJS/82/197 BURSTEIN D.: J/ApJS/94/687 Burstein D.: J/ApJS/96/123 J/MNRAS/292/499 J/MNRAS/305/259 J/MNRAS/321/277 VII/176A VII/198 Burton M.: J/AJ/130/1719 Burton M.G.: J/A+A/410/597 J/A+A/438/663 J/A+A/450/253 J/A+A/474/891 Burton M.G: J/MNRAS/291/261 Burton M.G.: J/MNRAS/301/640 J/MNRAS/363/405 J/MNRAS/367/1609 J/MNRAS/368/1223 Burton R.E.: J/AJ/131/84 Burton W.B.: J/A+A/391/159 J/A+AS/52/63 J/AJ/123/873 VIII/54 VIII/76 Burud I.: J/A+A/391/481 J/A+A/455/L1 Burwitz V.: J/A+A/458/541 J/A+A/465/375 Bus R.: J/AZh/72/22 BUS S.J.: J/AJ/107/1814

Busarello G.: J/A+A/389/787 J/A+A/397/431 J/A+A/424/79 J/A+A/425/797 J/A+A/428/1043

J/A+A/457/79 BUSARELLO G.: J/A+AS/105/433 Busarello G.: J/A+AS/127/117 J/A+AS/133/325 J/MNRAS/358/1116 BUSCEMA G.: J/AJ/100/1811 Buscher D.F.: J/AJ/114/2822 Buscombe W.: III/189A III/206 III/222 III/223 BUSCOMBE W.: III/52 Buser R.: J/A+AS/125/229 J/A+AS/130/65 BUSER R.: VI/78 Buser R.B.: VI/77 Busfield A.L.: J/A+A/461/11 J/A+A/474/891 J/MNRAS/366/1096 Bush T.C.: J/AJ/132/393 Bushimata T.: J/PASJ/52/631 Buson L.M.: J/A+AS/120/257 Busquets J.: J/A+AS/119/271 Bussons-Gordo J.: J/A+A/476/1191 Bustos G.: J/A+AS/136/1 BUTA N.S.: VI/15 VI/32 Buta R.: J/AJ/109/517 J/AJ/109/543 J/AJ/118/2071 J/AJ/121/225 J/ApJS/96/39 J/MNRAS/355/1251 Buta R.J.: J/AJ/108/2128 J/AJ/127/1982 J/MNRAS/274/1107 VII/137B VII/155 Butcher H.: J/ApJS/110/213 J/ApJS/122/51 Butenko G.Z.: J/other/KFNT/22.231 Butkovskaya V.: J/A+A/469/1069 Butler C.J.: J/A+AS/143/409

Butler K.: J/A+A/321/531 J/A+A/372/1078 J/A+A/372/1083 J/A+A/457/651 J/A+AS/108/1

J/A+AS/143/483 Butler N.: J/other/GCN/1070.1 Butler R.A.H.: J/ApJS/158/188 Butler R.P.: J/ApJ/634/625 J/ApJS/141/503 J/ApJS/152/261 J/MNRAS/370/163 J/MNRAS/372/163 Buttery H.: J/A+A/457/501 Buttery H.J.: VIII/70A VIII/78 VIII/81 Button S.: J/ApJS/45/97 BUTTRESS J.: J/MNRAS/250/314 Buxton M.M.: J/ApJ/608/L93 Buzasi D.L.: J/A+A/456/651 J/A+A/467/1215 BUZZONI A.: J/A+A/290/69 Buzzoni A.: J/A+A/456/985 J/A+A/464/201 J/A+A/471/127 J/A+A/473/429 BUZZONI A.: J/A+AS/103/245 Buzzoni A.: J/A+AS/119/391 J/MNRAS/246/576 J/MNRAS/375/1381 J/PASP/113/1365 Bychkov V.D.: J/A+A/407/631 J/A+A/430/1143 Bychkova L.V.: J/A+A/407/631 J/A+A/430/1143 Bychkova V. S.: J/AZh/81/925 Bykov O.P.: I/292 BYRAM E.T.: VII/80 Byrd G.G.: J/AJ/118/2071 J/AJ/127/1982 VI/76 BYRNE P.B.: J/A+AS/111/275 Byrne P.B.: J/A+AS/126/357 J/A+AS/126/81 J/A+AS/143/409 Bystrov N.F.: I/195 Byun Y.-I.: J/AJ/116/1775 J/AJ/117/2757 J/AJ/130/825 J/ApJ/619/L95 J/ApJS/164/38

C. H.: J/A+AS/135/145 C.: III/218 C. J.: J/PASP/108/682 C.: J/A+A/356/529 J/A+A/434/1125 J/A+A/465/271 J/A+A/475/915 J/A+AS/120/429 J/ApJ/527/573 Caballero J.A.: J/A+A/466/917 J/A+A/478/667 Cabanac R.: J/A+A/433/757 J/A+A/441/915 J/MNRAS/367/1478 Cabanac R.A.: J/ApJ/509/309 Cabanela J.E.: II/254 J/AJ/116/1094 J/ApJS/126/133 VII/214 CABANNE M.L.: J/PASP/98/1133 CABRERA-CANO J.: J/AJ/106/1547 Cabrera-Guerra F.: J/MNRAS/311/456 J/MNRAS/351/1290 Cabrera-Lavers A.: II/270 Cabrera-Lavers A.L.: J/A+A/413/1037 Cabrit S.: J/A+A/356/1010 CABRIT S.: J/A+AS/101/485 Cacciani A.: J/A+A/390/717 Caccianiga A.: J/A+A/428/383 J/A+A/476/1191 J/MNRAS/326/1455 J/MNRAS/329/877 J/MNRAS/348/937 CACCIARI C.: J/A+A/290/69 Cacciari C.: J/A+A/320/757 J/A+A/324/915 J/A+A/364/102 J/A+A/413/343 J/AJ/129/267 J/ApJ/640/L43

J/MNRAS/375/1381 CADONAU R.: II/189 Caffau E.: J/A+A/472/241 Cahill W.: IX/35 Cahn J.H.: J/A+AS/94/399

Caillault  J.-P.:  J/A+A/341/163 J/A+A/341/751 J/AJ/110/1229 J/ApJ/437/361 J/ApJ/445/280

J/ApJS/91/625 Cairos L.M.: J/A+A/429/115 J/ApJ/641/L13 Cakirli O.: J/A+A/388/298 J/A+A/409/959 Calabretta M.R.: J/AJ/128/16 Calamai G.: J/AJ/112/2786 Calbet X.: J/ApJ/491/L31 Caldwell J.A.R.: II/162 J/A+A/304/415 J/AJ/112/772 J/AJ/115/635 CALDWELL N.: J/AJ/106/986 Caldwell N.: J/AJ/112/2408 J/AJ/123/1216 J/AJ/127/1513 J/AJ/131/527 J/ApJ/619/193 J/ApJS/102/29

J/ApJS/126/271 J/ApJS/126/331 Cali A.: J/AJ/121/327

Caliskan H.: J/A+A/371/1078 J/A+A/394/187 J/A+A/406/975 J/A+A/447/685 J/A+AS/106/333

J/MNRAS/316/514 J/MNRAS/356/1256 J/PAZh/26/843 Calisse P.: J/MNRAS/367/1609 Calkins M.L.: J/AJ/131/527 J/ApJ/626/L5 Callas J.L.: J/ApJS/127/79 Caloi V.: J/A+A/464/953 Calura F.: J/A+A/445/93 Calvani M.: J/ApJS/145/199

Calvet N.: J/A+A/440/595 J/AJ/116/1816 J/AJ/128/805 J/AJ/129/363 J/AJ/129/856 J/AJ/129/907

J/AJ/130/188 J/ApJ/638/897 CALZETTI D.: III/167 Calzetti D.: J/AJ/110/2739 J/AJ/122/3046 J/ApJ/603/503 J/ApJ/612/481 J/ApJ/633/871 J/ApJS/143/377 Camargo I.B.: J/A+A/437/1135 Camargo J.I.B.: I/303 J/A+A/361/1143 J/A+A/375/308 J/A+A/383/1054 J/A+A/409/361 Cambresy L.: II/246 J/A+A/447/89 VII/237 Cambridge Astronomical Survey Unit: B/int Cambridge: I/304 Cameron A.: J/A+A/365/110 J/A+A/378/116 J/A+A/379/564

Cameron A.C.: III/153 J/MNRAS/358/795 J/MNRAS/362/117 J/MNRAS/372/1117 J/MNRAS/379/647

J/MNRAS/379/816 Cameron R.: J/ApJ/645/955 Cameron R.A.: J/ApJ/618/123 J/ApJS/150/19 CAMERON R.A.: VII/186

Camilo F.: J/ApJ/605/759 J/ApJ/615/897 J/ApJ/625/796 J/MNRAS/328/17 J/MNRAS/335/275

J/MNRAS/342/1299 J/MNRAS/352/1439 J/MNRAS/359/1524 J/MNRAS/372/777 VII/189 Campana S.: IX/34 J/A+A/435/799 J/A+A/448/L9 J/A+A/474/793 Campbell C.T.: VII/216 Campbell R.M.: J/A+A/472/763 Campusano L.E.: J/AJ/118/1468 J/AJ/118/1489 J/MNRAS/249/218 CAMPUSANO L.E.: J/MNRAS/266/317 Campusano L.E.: J/MNRAS/309/48 Canadian Astronomical Data-Center (CADC): B/cfht Canadian Astronomy Data Centre: B/jcmt Candia P.: J/ApJ/594/1 Canizares C.: J/ApJ/549/441 CANNON A.J.: III/1 Cannon A.J.: III/135A CANNON A.J.: III/138 III/155 III/182 III/99 Cannon R.: J/MNRAS/329/87 J/MNRAS/333/100 J/MNRAS/350/1485 J/MNRAS/352/939 J/MNRAS/372/425

J/MNRAS/375/68 CANNON R.D.: I/185 I/212 J/A+AS/111/75 Cannon R.D.: J/A+AS/122/507 J/A+AS/122/9 J/MNRAS/329/227 J/MNRAS/341/534 J/MNRAS/355/1196

J/MNRAS/362/689 Cannon W.: J/PASJ/52/997 Cano J.A.: J/A+AS/125/399 Canto Martins B.L.: J/A+A/458/895 Canzian B.: J/AJ/127/3034 J/AJ/127/3553 Cao A.: J/other/AcApS/16.103 VIII/44 Caon C.: J/ApJS/127/39 CAON N.: J/A+A/296/319 J/A+AS/106/199 Caon N.: J/ApJ/602/664 J/ApJ/641/L13 CAPACCIOLI M.: J/A+A/296/319 Capaccioli M.: J/A+A/333/479 J/A+A/389/787 J/A+A/428/339 CAPACCIOLI M.: J/A+AS/106/199 Capaccioli M.: J/A+AS/109/341 J/A+AS/114/527 J/MNRAS/358/1116 J/MNRAS/369/120 VII/3 VII/69 Capak P.: J/AJ/126/632 J/AJ/127/180 J/AJ/127/3121 J/AJ/127/3137 J/AJ/128/1483 Capalbi M.: J/A+A/448/L9 J/A+A/474/793 CAPDEVILLE A.: IV/16

Capelato H.V.: J/A+A/258/243 J/A+A/377/428 J/A+A/407/31 J/A+AS/114/565 J/ApJ/485/447

J/ApJS/110/1 CAPELATO H.V.: J/MNRAS/259/233 Caperton K.A.: J/AJ/130/2584 Capetti A.: J/A+A/439/487 Capilla G.: J/MNRAS/358/66 Capilla: J/A+A/394/479 Capitaine N.: J/A+A/355/398 J/A+A/400/1145 J/A+A/406/1135 J/A+A/450/855 J/A+A/459/981 Capitanio F.: J/ApJ/607/L33 J/ApJ/636/765 J/ApJS/170/175 Caplan J.: J/A+A/344/943 J/A+A/360/1107 J/A+A/394/225 J/A+A/399/1135 Capobianco C.C.: J/AJ/130/767 Cappellari M.: J/A+A/360/439 CAPPELLARO E.: II/103 Cappellaro E.: II/159 II/227 J/A+A/317/423 J/MNRAS/328/1181 J/MNRAS/369/1880 Cappellini B.: J/A+A/397/213 J/A+A/410/847 Cappelluti N.: J/A+A/462/449 Cappi A.: J/A+A/342/1 J/A+A/399/813 J/A+A/403/857 J/A+A/408/905 J/A+A/428/1043 J/A+A/430/19

J/A+A/457/79 J/A+A/461/81 J/A+AS/101/259 J/A+AS/130/323 Capria M.T.: J/A+A/461/789 Caputi K.I.: J/MNRAS/361/607 Caputo F.: J/A+AS/121/327 J/MNRAS/365/1357 Capuzzo-Dolcetta R.: J/A+A/357/437 CAPUZZO-DOLCETTA R.: J/A+AS/104/379 Caramazza M.: J/A+A/474/495 Carangelo N.: J/A+A/409/79 J/A+A/466/31 Caratti o Garatti A.: J/A+A/419/999 J/A+A/422/141 Caraveo P.: J/A+A/458/245 Carballo R.: J/A+AS/142/353 J/MNRAS/311/456 J/MNRAS/348/857 J/MNRAS/351/1290 J/MNRAS/353/211

J/MNRAS/370/1034 Carbillet M.: J/A+AS/134/545 Carbon D.F.: J/AJ/112/2274 Cardiel: J/MNRAS/326/959 Cardiel N.: J/A+AS/139/29 J/MNRAS/371/703 J/MNRAS/374/664 J/MNRAS/376/125 Cardini D.: J/A+A/374/1085 J/A+A/430/303 Cardona G.: J/ApJ/639/816 Cardona O.: J/A+A/347/583 J/AZh/76/83 Cardwell A.: J/A+A/367/498 J/A+A/414/23 Caretta C.A.: J/AJ/123/1200 J/AJ/128/2642 Carey L.: J/AJ/121/2308 J/AJ/121/2331 J/AJ/123/567 Carey S.J.: J/AJ/124/2990 J/AJ/126/1423 V/114 Carignan C.: J/A+A/420/147 J/A+A/466/905 Carilli C.: J/ApJ/606/664 Carilli C.L.: J/A+A/431/793 J/AJ/128/1974 Carini M.T.: J/AJ/121/90 Cariulo P.: VI/113 Carlberg R.G.: J/AJ/127/2455 J/AJ/130/2043 J/ApJ/471/694 J/ApJ/565/208 J/ApJS/102/289 J/ApJS/113/1

J/ApJS/116/211 J/ApJS/116/231 J/ApJS/116/247 J/ApJS/129/475 J/ApJS/157/228 Carleton N.P.: J/ApJ/620/961 Carlin J.L.: J/ApJ/617/262 Carlos De Breuck: VIII/69A CARLSBERG: I/126 Carlson B.: J/PASJ/52/997 Carlson G.: J/AJ/127/1131 Carlson M.N.: J/AJ/117/1700 CARLSTROEM U.: J/A+AS/93/121 CARNEY B.W.: J/A+A/291/121 Carney B.W.: J/A+A/415/521 J/A+A/415/531 CARNEY B.W.: J/AJ/107/2240 Carney B.W.: J/AJ/110/652 J/AJ/113/1073 J/AJ/117/2868 J/AJ/124/1144 J/AJ/129/267 J/AJ/129/656

J/AJ/130/1111 J/AJ/131/2274 J/AJ/131/431 CARNOCHAN A.: V/38 Carnochan D.J.: II/59B II/91 Caroli E.: J/ApJS/106/399 Carollo C.M.: J/AJ/119/486 J/ApJ/635/243 Carollo D.: J/A+A/448/579 Carpano S.: J/A+A/443/103 Carpenter J.: J/PASP/118/1690

Carpenter J.M.: II/246 J/AJ/114/198 J/AJ/121/3160 J/AJ/124/1001 J/AJ/129/1049 J/ApJ/540/236

J/ApJ/551/852 Carpenter K.G.: J/AJ/116/941 Carquillat J.M.: III/217 CARQUILLAT J.M.: IV/16 Carquillat J.-M.: J/A+A/354/881 Carquillat J.M.: J/A+AS/123/135 Carquillat J.-M.: J/A+AS/130/223 Carquillat J.M.: J/ApJS/143/513 CARQUILLAT J.M.: V/49 Carquillat J.M.: V/60 Carr B.J.: J/MNRAS/351/1071 Carr J.S.: J/ApJ/541/977 Carraminana A.: J/MNRAS/351/1290 VII/221 Carranza G.J.: J/AJ/128/62 CARRARO G.: J/A+A/300/128

Carraro G.: J/A+A/372/879 J/A+A/379/136 J/A+A/385/471 J/A+A/387/479 J/A+A/391/179 J/A+A/402/549 J/A+A/407/527 J/A+A/417/961 J/A+A/418/525 J/A+A/419/149 J/A+A/428/67 J/A+A/436/527 J/A+A/442/917 J/A+A/444/831 J/A+A/445/493 J/A+A/459/489 J/A+A/462/L9 J/A+A/464/573 J/A+A/466/181 J/A+A/466/931

CARRARO G.: J/A+AS/106/573

Carraro G.: J/A+AS/114/281 J/A+AS/121/451 J/AJ/130/652 J/MNRAS/346/18 J/MNRAS/364/179

J/MNRAS/367/1441 J/MNRAS/371/1301 J/MNRAS/374/372 Carrasco E.R.: J/AJ/121/148 J/ApJS/126/1 CARRASCO G.: I/43 Carrasco G.: I/82 CARRASCO G.: I/89A Carrasco G.: J/A+AS/100/489 CARRASCO G.: J/A+AS/105/429 Carrasco G.: J/A+AS/109/501 J/A+AS/126/311 CARRASCO G.: J/A+AS/95/355

Carrasco L.: J/A+A/349/619 J/A+A/376/775 J/A+A/396/449 J/A+A/401/471 J/A+A/422/925

J/A+A/425/429 J/A+A/446/839 J/A+AS/104/271 J/A+AS/112/257 J/AJ/124/1283 J/ApJ/453/616 Carreira P.: J/MNRAS/341/1066

Carrera F.J.: J/A+A/428/383 J/A+A/444/79 J/A+A/469/27 J/A+A/474/473 J/A+A/476/1191

J/MNRAS/311/456 J/MNRAS/318/1073 Carreta E.: J/A+A/464/927 J/A+A/464/967 J/A+AS/121/95

Carretta E.: J/A+A/314/191 J/A+A/320/757 J/A+A/350/955 J/A+A/354/169 J/A+A/356/238 J/A+A/404/187 J/A+A/413/343 J/A+A/420/683 J/A+A/430/603 J/A+A/440/901 J/A+A/455/271 J/A+A/464/939 J/A+A/464/953 J/AJ/110/2319 J/AJ/121/327 J/AJ/131/1766

Carretta: J/A+A/421/937

Carrier F.: J/A+A/341/469 J/A+A/385/488 J/A+A/394/151 J/A+A/398/1073 J/A+A/401/271

J/A+A/406/L23 J/A+A/470/1009 J/A+AS/130/223 Carrigan B.J.: J/AJ/116/2549 Carrigan Brian J.: J/AJ/113/401 Carrilli C.L.: VIII/52B CARRILLO D.: J/A+AS/113/483 Carrillo R.: J/AJ/126/1707 J/other/RMxAA/35.187 Carroll T.: J/MNRAS/373/781 Carruthers G.R.: II/92 CARRUTHERS G.R.: J/ApJ/408/484 Carruthers G.R.: J/ApJS/104/101 CARRUTHERS G.R.: J/ApJS/89/259 Carruthers G.R.: J/ApJS/96/605 Carson D.: J/MNRAS/372/425 Carson J.C.: J/ApJ/635/280 Carson R.: J/MNRAS/343/1131 Carswell R.F.: J/ApJ/636/631 J/MNRAS/335/555 CARTER B.: J/MNRAS/267/711 Carter B.D.: J/MNRAS/370/163 J/MNRAS/372/163 Carter B.J.: J/AJ/131/527 J/ApJ/559/606 Carter B.S.: J/MNRAS/321/77 Carter D.: J/AJ/123/2490 J/ApJ/274/534 J/ApJS/137/279 J/ApJS/72/715 J/MNRAS/369/120 J/MNRAS/369/97 Carter M. S.: J/AJ/110/880 Cartwright I.M.: J/A+AS/136/257 Cartwright J.: J/AZh/73/835 Casagrande L.: J/MNRAS/373/13 Casali M.M.: J/A+A/350/883 J/A+A/357/219 J/A+A/401/613 J/A+A/420/945 J/A+A/421/623 Casanova V.: J/A+A/449/937 J/A+A/462/591 Casas R.: J/A+AS/125/399 Casasola V.: J/A+A/422/941 Casassus S.: J/A+A/441/181 Caselli P.: J/A+A/413/1177 J/A+AS/103/541 J/MNRAS/276/57 Caserini E.: J/A+A/410/847

Casertano S.: J/A+AS/137/101 J/AJ/120/2747 J/ApJ/600/L155 J/ApJ/607/665 J/ApJ/613/200

J/PASP/108/682 Casewell S.L.: J/MNRAS/365/447 Casey B.: J/A+A/304/415 Casey B.W.: J/A+AS/128/251 J/A+AS/130/245 Casian B.: I/284 Casini R.: J/A+AS/114/363 Casoli F.: J/A+A/327/522 J/A+A/411/381 J/A+AS/116/193 Casperson D.: J/AJ/119/1901 J/AJ/127/2436 Cass C.P.: J/MNRAS/362/689 Cassar L.: I/123 CASSAR L.: I/142 Cassata P.: J/A+A/465/1099 Cassatella A.: J/A+A/317/712 J/A+A/373/730 J/A+A/374/1085 J/A+A/420/571 CASSATELLA A.: J/ApJS/71/983 Cassatella A.: J/ApJS/99/223 Cassinelli J.P.: J/A+AS/118/481 Cassisi S.: J/A+A/380/478 J/A+A/432/851 J/A+A/464/939 J/A+A/464/953 J/A+A/464/967 Castander F.: J/MNRAS/311/456 Castander F.J.: J/AJ/119/2598 J/AJ/120/1 J/AJ/121/2308 J/AJ/121/2331 J/AJ/123/567 J/AJ/125/1817

J/AJ/126/2579 J/AJ/134/1103 J/ApJ/621/104 J/ApJS/162/1 J/PASJ/55/771 Castaneda H.O.: J/AJ/110/212 Castelaz M.: II/95 Castellani M.: J/A+A/310/715 J/A+A/330/453 J/AJ/118/1671 Castellani V.: J/A+A/351/526 J/A+AS/121/327 CASTELLANI V.: J/AJ/102/137 Castellani V.: J/ApJS/71/25 VI/113 Castelli F.: III/238 CASTELLI F.: J/A+A/281/817 J/A+A/290/885

Castelli F.: J/A+A/314/191 J/A+A/319/593 J/A+A/333/231 J/A+A/364/102 J/A+A/391/1039

J/A+A/417/1055 J/A+A/425/263 J/A+A/454/333 J/ApJ/634/1319 J/MNRAS/371/1503 Castermans T.: J/other/APh/26.282 Castets A.: J/A+A/309/581 J/A+A/417/107 Castilho B.: VI/120 CASTILHO B.V.: J/A+A/297/503 Castilho B.V.: J/A+A/364/674 Castro Ceron J.M.: J/A+A/427/87 CASTRO S.: J/A+AS/111/17 Castro S.: J/ApJ/612/447 Castro-Rodriguez N.: J/A+A/391/519 J/A+A/399/457 Castro-Tirado A.J.: J/A+AS/129/1 J/ApJ/641/L13 Casuso E.: J/ApJS/111/203 Caswell J.L.: J/A+A/171/261 J/A+AS/137/43 J/A+AS/90/327 VII/14 VIII/20 Catalano F.A.: III/152 III/162A III/199A J/A+A/378/113 CATALANO F.A.: J/A+AS/108/595 Catalano F.A.: J/A+AS/121/57 Catalano S.: J/A+A/333/205 J/A+A/388/298 J/AN/328/938 Catanzaro G.: J/A+A/464/939 J/MNRAS/368/247 Catarzi M.: J/A+AS/103/541 J/A+AS/134/115 CATCHPOLE R.: J/MNRAS/267/711 Catelan M.: J/A+A/474/121 J/AJ/122/2569 J/ApJ/640/L43 J/PASP/117/1325 Catinella B.: J/AJ/130/1037 J/AJ/130/2613 J/AJ/131/716 Caux E.: J/A+A/330/990 J/A+A/425/927 Cavalier F.: J/A+A/303/137 CAVALIER F.: J/A+AS/109/447 Cavallo R.M.: J/AJ/120/1364 Cavallone M.: J/A+AS/130/109 J/A+AS/132/361 Cavazzuti E.: J/A+A/472/699 CAWSON M.: J/AJ/100/1091 Cawthorne T.V.: J/AJ/130/1418 Cay M.T.: J/A+A/394/187 Cayatte V.: J/A+A/342/1 J/A+A/415/9 J/A+A/464/L21 CAYATTE V.: J/A+AS/113/35 Cayatte V.: J/A+AS/130/323 J/ApJ/520/491 CAYREL de STROBEL G.: III/165 Cayrel de Strobel G.: III/200B III/221A CAYREL de STROBEL G.: III/96 Cayrel de Strobel G.: J/A+A/331/81 Cayrel R.: J/A+A/403/1105 J/A+A/404/661 J/A+A/416/1117 J/A+AS/133/221 Cazzoli G.: J/A+A/413/1177 C.B.: III/168 C.E.: V/111 Ceballos M.T.: J/A+A/444/79 J/A+A/469/27 J/A+A/476/1191 Cebral J.R.: J/A+AS/109/201 Ceccarelli C.: J/A+A/472/1029 Cederblad S.: VII/231 Cedres B.: J/A+A/391/809 J/A+A/422/511 Cellone S.A.: J/A+A/409/857 Celotti A.: J/MNRAS/341/993 J/MNRAS/357/1267 Cen R.: J/ApJS/163/80 Cen Xuefen: J/A+AS/114/337 J/A+AS/128/315 J/A+AS/129/569 Cenarro A.J.: J/AJ/131/814 J/MNRAS/371/703 J/MNRAS/374/664 Cenarro: J/MNRAS/326/959 Cenci A.: J/A+AS/125/459 Centonza M.: J/A+A/363/887 Cepa J.: J/A+A/391/809 J/A+A/422/511 J/A+AS/128/545 CEPA J.: J/A+AS/93/557 J/AJ/106/56 Cepa J.: J/AJ/110/212 J/AJ/122/38 Ceriani L.: J/ApJ/543/178 Cerniauskas A.: J/BaltA/12/323 Cernicharo J.: J/A+A/329/219 J/A+A/451/1125 J/A+A/466/255 J/A+AS/142/181 CERNIS K.: II/196

Cernis K.: II/197 J/BaltA/10/319 J/BaltA/10/349 J/BaltA/11/219 J/BaltA/11/231 J/BaltA/11/321

J/BaltA/11/417 J/BaltA/12/301 J/BaltA/12/323 J/BaltA/13/555 Cerqueira A.H.: J/AJ/130/1707 Cersosimo J.C.: J/A+AS/97/435 Cersosimo S.: J/AJ/112/2408 Cervino M.: J/MNRAS/357/945

Cesaroni R.: J/A+A/308/573 J/A+A/336/339 J/A+A/368/845 J/A+A/417/615 J/A+A/432/921

J/A+A/447/221 J/A+A/453/911 J/A+AS/101/153 J/A+AS/103/541 Cesarsky C.: J/A+A/357/219 J/A+A/372/364 J/A+A/382/60 J/A+A/427/773 J/A+A/430/115

J/A+A/434/53 J/A+A/454/423 J/A+A/465/1099 J/A+A/478/83 J/ApJ/517/148 J/ApJ/552/289 J/MNRAS/302/222 J/MNRAS/305/297

Cesarsky C. Perez-Fournon I.: J/A+A/427/23

Cesarsky C.J.: J/A+A/342/313 J/A+A/350/883 J/A+A/372/173 J/A+A/420/945 J/A+A/421/623

J/A+A/423/867 J/MNRAS/351/1290 Cesarsky D.: J/A+A/372/173 CHABOD D.: I/74 Chaboyer B.: J/AJ/112/2408 J/AJ/122/3219 J/AJ/131/1565 J/AJ/131/2538 Chabrier G.: J/A+A/327/1039 J/A+A/328/83 J/A+A/337/403 Chadid M.: J/A+A/308/481 Chadima P.: J/A+A/463/233 Chaffee F.: J/A+A/442/109 Chaffee F.H.: J/AJ/109/1498 J/AJ/122/518 J/AJ/123/2223 J/AJ/127/3121 J/ApJS/126/133 J/ApJS/135/227 Chagnon G.: J/A+A/393/183 CHALABAEV A.: J/A+A/285/1

Challis P.: J/A+A/428/555 J/AJ/130/2453 J/AJ/131/1648 J/AJ/131/527 J/ApJ/594/1 J/ApJ/602/571 J/ApJ/607/665 J/ApJ/613/200 J/ApJ/626/L5 J/ApJS/99/223 Challis P.M.: J/ApJ/627/579 Chalonge D.: III/187 CHAMARAUX P.: J/A+A/299/347 Chamaraux P.: J/A+AS/104/529 J/AJ/117/2039 J/MNRAS/307/236 Chamberlin R.A.: J/ApJ/517/282 Chambers E.T.: J/ApJ/639/227 Chambers K.C.: J/A+AS/108/79 J/ApJ/602/571 J/ApJS/109/367 Chambliss C.R.: J/AJ/106/2058 Champeaux J.-P.: J/A+A/439/387 Chan B.: J/ApJS/155/1 Chan B.H.P.: J/MNRAS/352/1245 CHAN K.-W.: VI/56 Chan S.J.: J/A+AS/115/285 Chaname J.: J/ApJS/150/455 Chandar R.: J/A+A/366/498 J/AJ/122/1330 J/ApJS/122/431 Chandler C.J.: J/A+A/440/151 J/ApJ/628/789 J/MNRAS/354/485

Chandra S.: II/40 J/A+A/376/333 J/A+A/376/356 J/A+A/402/1 J/A+A/423/235 J/A+AS/114/175

J/A+AS/117/557 J/A+AS/131/137 J/A+AS/142/113 Chandra Team: B/chandra Chandrasekhar T.: J/AJ/116/1447 Chanial P.: J/A+A/382/60 J/A+A/462/81 Chapellier E.: J/A+A/304/406 J/A+A/426/247 J/A+A/448/697 Chapin E.L.: J/MNRAS/372/1621 CHAPMAN J.: J/MNRAS/267/711 Chapman J.M.: J/A+A/366/481 J/A+AS/119/459 J/A+AS/122/79 J/A+AS/124/509 J/ApJS/155/595 Chapman S.: J/A+A/424/371 J/MNRAS/372/1621 J/MNRAS/375/831 Chapman S.C.: J/ApJ/614/671 J/ApJ/616/71 J/ApJ/622/772 J/MNRAS/380/281 Chapront J.: J/A+A/311/710 J/A+A/400/1145 CHAPRONT J.: VI/66 VI/79 Chapront J.: VI/87 VI/88 CHAPRONT-TOUZE M.: VI/66 VI/79 Char S.: J/A+A/389/149 Charbonel C.: J/A+AS/115/339 Charbonneau D.: J/A+A/476/1347 Charbonnel C.: J/A+A/442/961 CHARBONNEL C.: J/A+AS/101/415 J/A+AS/102/339 J/A+AS/103/97 Charbonnel C.: J/A+AS/135/405 J/AJ/127/373 Charbonnel C. Maeder A.: J/A+AS/128/471 Chareton M.: I/196 J/A+A/374/132 Charitonova G.A.: J/A+AS/139/47 Charles E. Worley: I/281 I/301 Charles P.A.: J/ApJ/636/1002 J/MNRAS/274/1194 Charlot P.: J/A+A/437/1135 J/AJ/127/3587 J/AJ/130/2529 J/ApJS/128/17 Charlot S.: J/A+A/396/503 J/A+A/428/1043 J/A+A/444/365 J/A+A/448/893 J/A+A/457/79 Charlot X.: J/A+A/389/149 J/A+A/454/185 Charlton J.C.: J/ApJ/613/129

Charmandaris V.: J/A+A/347/841 J/A+A/381/941 J/A+A/422/129 J/ApJ/599/1049 J/ApJ/634/128

J/ApJ/634/L1 Charpinet S.: J/A+A/459/557 Charro E.: J/A+A/376/1106 J/A+A/387/1146 Chartas G.: J/AJ/122/2810 J/AJ/126/539 J/ApJ/554/742 Chary R.: J/ApJ/634/128 Chary R.R.: J/A+A/451/57 Chary R.-R.: J/ApJ/640/603 Chastenay P.: J/MNRAS/246/169 Chauvet F.: J/A+AS/125/399 Chauville J.: J/A+A/378/861 Chavarria C.: J/ApJS/117/319 CHAVARRIA-K C.: J/A+AS/104/557 Chavarria-K C.: J/A+AS/145/187 J/other/RMxAA/38.141 Chaves O.L.: J/AJ/116/1 J/AJ/118/1131 J/AJ/126/2268 Chaves T.: J/ApJ/598/969 J/MNRAS/371/147 Chavez M.: J/A+AS/126/267 J/ApJ/601/485 J/ApJ/613/312 J/ApJ/634/1319 J/PASP/113/1365 Chavira E.: II/131 J/AZh/76/83 J/AZh/77/3 Chavushyan V.: J/A+A/414/487 J/ApJ/639/37 Chavushyan V.H.: J/A+A/376/775 J/A+A/422/925 J/A+AS/147/169 J/AJ/124/1283 J/ApJ/523/87 Chayer P.: J/A+A/459/557 Chelli A.: J/A+A/349/619 Chelovekov I.V.: J/PAZh/32/508 CHEMESE M.: VI/60 Chemin L.: J/A+A/464/L21 Chen A.: J/MNRAS/287/867 Chen A.B.-C.: J/AJ/126/762 Chen A.W.: J/ApJS/123/79 Chen B.: J/AJ/124/2164 Chen C.W.: J/BaltA/13/555 Chen F.Z.: J/AJ/112/628 Chen G.X.: J/A+A/408/789 J/A+AS/136/395 Chen G.-X.: J/A+AS/147/111 Chen H.: J/A+AS/99/545 J/ApJ/540/1016 J/ApJ/566/993 VIII/44 Chen H.-W.: J/AJ/123/2206 J/AJ/127/2455 J/AJ/130/1324 J/ApJ/498/77 J/ApJ/538/493 J/ApJ/559/654

J/ApJ/617/718 J/ApJ/643/680 J/ApJS/135/41 Chen J.: J/A+A/397/361 J/A+A/449/143 J/AJ/117/2757 J/AJ/122/1718 J/AJ/126/1286 J/AJ/127/2579

J/AJ/129/1818 J/AJ/131/2245 VII/168 Chen J.S.: J/AJ/112/628 Chen J.-S.: J/PASP/118/1104 Chen L.: I/257B I/258 J/A+AS/142/373 J/A+AS/145/223 J/AJ/125/1397 J/other/ChJAA/6.435 Chen M.Z.: J/MNRAS/328/855 Chen P.S.: J/A+A/368/1006 J/A+A/372/245 Chen P.-S.: J/A+A/387/129

Chen P.S.: J/A+A/434/201 J/A+AS/113/51 J/AJ/131/1942 J/other/NewA/10.325 J/other/NewA/11.27

J/other/NewA/11.557 J/other/NewA/9.509 Chen R.: J/A+AS/126/503 J/AJ/117/2757 Chen W.P.: I/123 CHEN W.P.: I/142 Chen W-P.: J/AJ/112/628 Chen W.-P.: J/AJ/117/2757 J/AJ/126/1286 Chen W.P.: J/BaltA/13/555 Chen X.: J/ApJ/640/982 Chen Y.: J/AJ/117/446 J/AJ/123/578 J/ApJ/629/288 J/other/AcApS/20.366 Chen Y.Q.: J/A+A/371/943 J/A+A/397/257 J/A+A/415/993 J/A+A/425/671 Chen Y.Q: J/A+AS/141/491 Chen Y.Q.: J/ApJS/126/461 J/MNRAS/370/2091 Chen Y.-Q.: J/PASJ/58/389 Cheng E.S.: J/AJ/128/1990 J/ApJ/612/202 J/ApJ/621/53 J/ApJS/150/1 Cheng F.Z.: J/A+A/389/845 J/A+A/396/503 Cheng F.-Z.: J/ApJ/510/659 Cheng H.: J/A+AS/132/233 Cheng K.-P.: J/AJ/108/1350 J/AJ/112/2408 J/AJ/112/601 J/AJ/113/1011 CHENG K.-P.: J/ApJ/413/604 Cheng K.-P.: J/ApJ/438/181 J/ApJ/446/622 Cheng K.P.: J/ApJS/98/595 Cheng K.S.: J/MNRAS/328/855 Cheng X.: J/MNRAS/376/1707 Chengalur J.N.: J/ApJS/89/189

Chentsov E.L.: J/A+A/397/1035 J/A+AS/131/469 J/A+AS/147/5 J/AZh/81/1104 J/AZh/81/333

J/PAZh/32/737 Chepurnov A.V.: J/A+AS/101/447 Cherchneff I.: J/A+A/330/676 J/A+A/357/572 Cherepashchuk A.M.: J/PAZh/30/430 J/PAZh/30/589 Chernin A.D.: J/AZh/78/963 Chesneau O.: J/A+A/465/469 J/A+A/466/L1 Chester M.: J/A+AS/134/221 Chester T.: II/156A II/246 J/ApJ/560/566 Cheung A.S.-C.: VI/61 Chevallier L.: J/A+A/429/235 Chevreton M.: J/A+A/468/637 Chiaberge M.: J/A+AS/121/119 J/A+AS/127/445 J/A+AS/129/577 J/A+AS/130/109 Chiang J.: J/ApJS/94/551 Chiappetti L.: J/A+A/439/413 J/A+A/467/73 J/A+A/474/473 J/MNRAS/382/279 Chiappini C.: J/A+A/425/489 J/MNRAS/336/879 Chiba M.: J/A+A/397/899 J/AJ/115/168 J/AJ/119/2866 Chiboucas K.: J/AJ/132/347 J/ApJS/170/95 Chidichimo M.C.: J/A+A/365/266 J/A+A/430/331 J/A+A/432/1137 Chieffi A.: J/AJ/118/2839 J/ApJ/608/405 Chies-Santos A.L.: J/A+A/455/453 J/A+A/467/1003 Chihara H.: J/A+A/449/583 Chilingarian I.: J/A+A/422/925 Chin Y.-N.: J/A+A/305/960 Chincarini G.: J/A+A/342/1 J/A+A/425/367 J/A+A/448/L9 J/A+A/469/503 J/A+A/474/793 CHINCARINI G.: J/A+AS/103/245 J/A+AS/110/279

Chincarini G.: J/A+AS/119/391 J/A+AS/130/323 J/A+AS/137/101 J/A+AS/137/495 J/A+AS/140/155

J/A+AS/145/111 J/ApJ/514/148 J/MNRAS/246/576 Chini R.: J/A+A/376/1123 J/A+A/426/97 J/A+A/456/953 J/A+A/459/55 J/A+A/473/369 J/A+AS/118/47

Chiosi C.: II/82 J/A+A/240/262 J/A+A/311/361 J/A+A/335/823 J/A+A/345/419 J/A+A/361/1023

J/A+A/392/1129 J/A+AS/106/275 CHIOSI C.: J/A+AS/108/571 Chiosi C.: J/A+AS/130/251 J/A+AS/141/371 J/AJ/110/212 VI/96 Chiosi E.: J/A+A/452/179 Chirkin D.: J/other/APh/26.282 Chis D.: J/A+AS/125/399 Chisholm J.R.: J/ApJ/596/437 Chiu K.: J/AJ/127/3553 J/AJ/131/2722 J/MNRAS/372/425 Chiueh T-H.: J/AJ/112/628 Chiumiento G.: J/A+AS/121/119 Chlevicki G.: III/130 Cho S.H.: J/ApJS/106/463 Cho S.-H.: J/ApJS/165/360 Choi P.I.: J/AJ/131/250 Choi Y.J.: J/ApJ/525/420 Chomiuk L.: J/AJ/127/1943 J/AJ/130/2571 Chomiuk L.B.: J/AJ/130/2584 Chong S.S.: J/ApJS/155/257 Chornock R.: J/AJ/130/2453 J/AJ/131/1648 J/ApJ/600/L99 J/ApJ/602/571 J/ApJ/607/665 J/ApJ/613/200

J/ApJ/627/579 CHRISTENSEN C.G.: J/AJ/106/2493 Christensen J.: J/AJ/119/486 J/ApJ/551/973 CHRISTENSEN P.R.: J/A+AS/109/171 Christensen P.R.: J/AJ/117/2329 Christian C.: J/A+A/235/185 J/ApJS/102/29 Christian C.A.: J/AJ/107/751 J/PASP/109/998 Christian D.: J/ApJS/96/303

Christian D.J.: IX/35 J/A+A/384/890 J/A+A/467/785 J/MNRAS/372/1117 J/MNRAS/379/647

J/MNRAS/379/816 Christiansen W.A.: J/AJ/123/1216 J/AJ/130/957 J/AJ/131/1280 Christlein D.: J/AJ/134/1103 J/ApJ/591/764 J/ApJS/162/1

Christlieb N.: J/A+A/358/77 J/A+A/363/1040 J/A+A/378/556 J/A+A/430/223 J/A+A/431/143

J/A+A/439/129 J/AJ/132/137 J/ApJS/161/147 Christlieb NN.: J/A+A/375/366 Christodoulou D.M.: J/PASP/108/772 Christopher M.: J/AJ/126/1916 Christopher N.M.: J/MNRAS/365/439 CHROMEY F.: J/AJ/107/1453 Chrysostomou: J/A+A/385/1014 Chu Y.-H.: J/A+A/418/841 J/AJ/125/2239 J/AJ/127/477 J/ApJ/562/446 Chu Y.Q.: J/AJ/112/628 CHUBB T.A.: VII/80 Chugainov P. V.: J/A+A/323/113 CHUGAINOV P.F.: J/A+AS/103/183 Chun M.-S.: J/A+A/437/61 Chun M.S.: J/A+A/445/69 Chun M.-S.: J/A+A/454/717 J/AJ/116/1775 Chun M.-Y.: J/AJ/120/333 J/AJ/128/1684 Chung H.-S.: J/ApJS/165/360 Chupp E. L.: J/ApJS/120/409 Churazov E.: J/A+A/475/775 J/ApJ/562/254 Churazov E.M.: J/PAZh/30/430 Church M.J.: J/MNRAS/344/134 Churchwell E.: J/A+AS/120/283 J/ApJS/154/553 J/ApJS/91/659 Churchwell E.B.: J/A+A/305/960 J/ApJ/635/560 Chuvaev K. K.: J/AZh/71/200 Chuvaev K.K.: J/A+A/320/405 J/PAZh/23/409 Chyzy K.T.: J/A+A/303/420 Ciardi D.R.: J/PASP/114/656 Ciardullo R.: J/A+A/306/9 J/AJ/99/1079 J/ApJ/339/53 J/ApJ/344/704 J/ApJ/344/715 J/ApJ/356/332

J/ApJ/365/471 J/ApJ/383/487 J/ApJ/414/454 J/ApJ/414/463 J/ApJ/416/62 J/ApJ/458/455 J/ApJ/462/1 J/ApJ/479/231 J/ApJ/492/62 J/ApJ/577/31 J/ApJ/582/170 J/ApJ/614/167 J/ApJS/145/65 J/ApJS/88/423

CIATTI F.: II/103 Cicek C.: J/A+A/387/240 J/A+A/417/263 Cidale L.: J/A+A/378/861 Cidale L.S.: J/A+A/385/986 Cieslinski D.: J/A+AS/124/55 J/A+AS/131/119 Cignoni M.: J/MNRAS/371/1503 Ciliegi P.: J/A+A/372/364 J/A+A/398/901 J/A+A/403/857 J/A+A/409/79 J/A+A/428/1043 J/A+A/432/69

J/A+A/457/501 J/A+A/457/79 J/A+A/463/519 J/A+A/466/31 J/AJ/128/1974 J/MNRAS/302/222 J/MNRAS/305/297 J/MNRAS/325/1173 J/MNRAS/342/575 J/MNRAS/351/1290 J/MNRAS/358/333 J/MNRAS/358/397

Cimatti A.: II/253 J/A+A/392/395 J/A+A/432/69 J/A+A/437/883 J/A+A/465/1099 Cimo G.: J/A+A/402/151 Cincunegui C.: J/A+A/414/699 J/A+A/469/309 Cinzano P.: J/A+A/342/671 Cioni M.-R.: II/228A

Cioni M.-R.L.: J/A+A/377/945 J/A+A/406/51 J/A+A/429/837 J/A+A/438/521 J/A+A/442/165

J/MNRAS/376/313 Ciprini S.: J/A+A/395/11 J/A+A/402/151 J/A+A/469/503 Cirasuolo M.: J/MNRAS/341/993 J/MNRAS/357/1267 Cirimele G.: J/A+AS/110/313 J/A+AS/125/459 J/A+AS/146/373 C.J.: J/A+A/385/1014 Claes P.: J/A+A/357/839 CLAIREMIDI J.: I/74 Clampin M.: J/AJ/128/1990 J/ApJ/612/202 J/ApJ/612/481 J/ApJ/621/53 J/ApJS/150/1 Clampitt L.: J/AJ/114/699

Claret A.: J/A+A/335/647 J/A+A/336/920 J/A+A/340/196 J/A+A/346/586 J/A+A/363/1081 J/A+A/372/588 J/A+A/401/657 J/A+A/424/919 J/A+A/428/1001 J/A+A/440/647 J/A+A/443/157 J/A+A/453/769 J/A+A/454/171 J/A+A/454/179 J/A+A/459/871 J/A+A/462/591 J/A+A/467/1389 J/A+A/469/285

CLARET A.: J/A+AS/110/329 Claret A.: J/A+AS/114/247 J/A+AS/131/395 J/AJ/119/1389 J/AJ/128/1324 J/AJ/128/1340 J/AJ/128/1840

J/AJ/130/2838 J/MNRAS/362/1025 J/MNRAS/363/529 VI/118 Claria J.J.: J/A+A/324/91 J/A+A/360/529 J/A+A/361/945 J/A+A/370/931 J/A+A/375/30 J/A+A/379/453

J/A+A/388/179 J/A+A/399/543 J/A+A/409/541 J/A+A/418/979 J/A+A/421/991 J/A+A/423/189 J/A+A/440/111 J/A+AS/107/39 J/A+AS/127/423 J/A+AS/128/131 J/A+AS/134/301 J/AJ/116/723 J/ApJS/116/263 J/MNRAS/335/233 J/MNRAS/340/1249 J/MNRAS/346/390 J/MNRAS/349/641 J/MNRAS/367/599 J/MNRAS/377/1737 J/other/NewA/11.262 J/PASP/117/22

Clark B.G.: J/AJ/98/419 Clark D.: J/MNRAS/358/1215 J/MNRAS/377/300 Clark D.H.: VII/14 Clark D.J.: J/ApJS/170/175 Clark D.M.: J/ApJ/635/280 Clark F.O.: J/ApJ/488/224 Clark L.L.: J/AJ/128/3019 J/ApJ/565/511 Clark P.J.: J/A+A/459/557 Clarke A.J.: J/A+A/461/11 J/MNRAS/363/1111 Clarke C.: J/MNRAS/370/954 J/MNRAS/377/741 Clarke D.: J/A+A/383/580 Clarke J.: J/AJ/113/1001 Clarke J. T.: J/PASP/108/682 Clarke J.T.: J/AJ/117/1700 Clarkson W.I.: J/A+A/467/785 J/MNRAS/372/1117 J/MNRAS/379/647 J/MNRAS/379/816 Classen H.-T.: J/A+AS/119/489 Claudi R.: J/A+A/404/187 J/A+A/464/939 J/A+A/464/953 J/A+A/471/515 J/AJ/121/327 Claudi R.U.: J/A+A/377/123 J/A+A/420/683 J/A+A/454/581 Claus F.: J/other/IBVS/5495

Clausen J.V.: J/A+A/352/605 J/A+A/364/455 J/A+A/392/215 J/A+A/402/509 J/A+A/456/651

J/A+A/461/1065 J/A+A/461/1077 CLAUSEN J.V.: J/A+AS/101/563 Clausen J.V.: J/A+AS/109/425 CLAUSEN J.V.: J/A+AS/112/123

Clausen J.V.: J/A+AS/115/315 J/A+AS/122/559 J/A+AS/123/1 J/A+AS/123/59 J/A+AS/125/471

J/A+AS/128/251 J/A+AS/130/245 J/A+AS/146/365 Clavel J.: J/A+A/357/839 J/A+A/369/57 J/A+AS/121/461 Claver C.F.: J/AJ/128/1825 Claver J.: J/AJ/123/3154 Clayton G.C.: J/ApJS/143/377 Cleary K.: J/MNRAS/341/1066 Cleary K.A.: J/MNRAS/360/340 Cleary M.N.: VIII/8A VIII/8A Clegg R.E.S.: J/A+AS/135/359 J/MNRAS/234/583 Clem J.: J/other/APh/26.282 Clem J.L.: J/AJ/127/1227 CLEMENS C.M.: VII/144 Clemens C.M.: VII/164 VII/193 Clemens D.P.: J/ApJ/582/756 J/ApJ/635/560 J/ApJS/63/821 VII/114 Clement C.: J/ApJ/542/257 V/97 V/97 Clement C.M.: J/AJ/113/1711 J/AJ/114/1190 J/AJ/122/2587 J/AJ/127/334 CLEMENT C.M.: J/ApJ/412/183 Clement R.: J/A+A/304/415 J/A+AS/123/1 J/A+AS/123/59 J/A+AS/128/139 Clementini G.: J/A+A/430/603 J/A+A/431/1189 CLEMENTINI G.: J/A+AS/84/59 Clementini G.: J/AJ/110/2319 CLEMENTINI G.: J/MNRAS/252/357 Clementini G.: J/MNRAS/264/273 J/MNRAS/363/734 Clementini: J/A+A/421/937 Clements D.L.: J/A+A/372/364 J/MNRAS/279/459 J/MNRAS/351/1290 J/MNRAS/372/1621 Clements S.D.: J/AJ/121/90 Clenet Y.: J/A+A/466/949 Clewley L.: J/MNRAS/352/285 J/MNRAS/379/1343 Cline T.: J/ApJS/120/399 J/ApJS/156/217 J/other/GCN/1070.1 Clocchiatti A.: J/AJ/130/2453 J/AJ/131/1648 J/ApJ/594/1 J/ApJ/602/571 Close L.M.: J/AJ/117/1375 Clowe D.: J/MNRAS/371/1777 Clowe D.I.: J/A+A/427/397 J/A+A/444/365 Clowe: J/A+A/394/395 CLOWES R.: J/A+AS/112/347 Clowes R.: J/A+AS/119/265 Clowes R.G.: J/MNRAS/249/218 CLOWES R.G.: J/MNRAS/266/317 Clowes R.G.: J/MNRAS/309/48 J/MNRAS/369/1334 C.M.: J/A+A/400/533 C.N.R.S.(Paris): I/29 Coarasa J.A.: J/A+A/410/813 Coarasa T.: J/A+A/406/L9 Cobb B.E.: J/ApJ/608/L93 Coccato L.: J/A+A/408/873 J/A+A/416/507 J/A+A/440/107 J/MNRAS/369/120 Cocchi M.: J/ApJ/607/L33 Cocchia F.: J/A+A/409/79 J/A+A/466/31 J/AJ/127/3075 Cochran W.D.: J/AJ/127/3579 Code A.D.: II/67 II/83 VI/105 CODELLA C.: J/A+A/284/233 J/A+A/291/261 Codella C.: J/A+A/327/1177 J/A+A/327/755 J/A+A/417/615 J/A+AS/103/541 J/MNRAS/276/57 Coe D.: J/ApJ/621/53 Coelho P.: J/A+A/404/661 J/AJ/130/55 VI/120 Coenda V.: J/AJ/129/1237 Cofino A.S.: J/MNRAS/353/211 Cohen A.S.: J/A+A/456/791 J/A+A/471/1105 J/ApJ/591/640 J/ApJS/150/417 J/MNRAS/355/20 VIII/79 Cohen J.G.: J/A+A/391/481 J/AJ/112/1489 J/AJ/120/2190 J/AJ/121/2895 J/AJ/122/611 J/AJ/127/1545

J/AJ/127/1579 J/AJ/129/1607 J/AJ/129/303 J/AJ/130/1177 J/AJ/132/137 J/ApJ/336/780 J/ApJ/471/L5 J/ApJ/538/29 J/ApJ/591/850 J/ApJS/127/1

COHEN M.: J/AJ/105/1860 Cohen M.: J/AJ/110/275 J/AJ/112/2274 J/AJ/112/241 J/ApJ/635/560 J/MNRAS/362/689 J/PASP/118/1690

V/107 V/114 V/127 Cohen M.H.: J/AJ/130/2473 J/ApJ/430/467 J/ApJ/609/539 COHEN R.D.: J/AJ/106/426 Cohen R.E.: J/AJ/127/1602 Cohen R.J.: J/A+A/325/255 Cohen R.L.: J/AJ/113/669 Cohen R.S.: VIII/39 Cohen S.H.: J/ApJ/639/724 Cohn H. N.: J/AJ/115/708 Cohn H.N.: J/AJ/123/2261 J/ApJ/625/796 Coia D.: J/A+AS/141/257 Coil A.: J/ApJ/633/174 Coil A.L.: J/AJ/127/3121 J/ApJ/594/1 J/ApJ/625/6 J/ApJ/634/977 J/MNRAS/371/221 Colafrancesco S.: J/A+AS/133/403 Colas F.: J/A+AS/125/399 Colavita M.M.: J/AJ/130/2866 J/ApJ/636/1020 Colbert E.J.M.: J/ApJ/602/231 J/ApJS/143/25 Colbert J.W.: J/ApJ/634/128 Coldwell G.V.: J/ApJS/155/33 Cole A.A.: J/AJ/120/1808 J/AJ/123/3154 J/AJ/125/1261 J/AJ/129/1465 J/MNRAS/327/918 Cole S.: J/MNRAS/329/227 J/MNRAS/329/87 J/MNRAS/333/100 J/MNRAS/350/1485 J/MNRAS/352/939 Coleman M.G.: J/AJ/130/1065 J/AJ/131/912 Coleman P.H.: J/AJ/90/1437 J/ApJS/107/239 Colin J.: I/272 I/303 COLIN J.: I/74 Colin J.: J/A+AS/119/307 J/A+AS/125/399 COLINA L.: J/ApJS/91/507 Colla: VIII/36 Colless M.: J/ApJ/499/112 J/ApJS/106/1 J/MNRAS/224/453 COLLESS M.: J/MNRAS/244/408 J/MNRAS/253/686 Colless M.: J/MNRAS/273/157 J/MNRAS/305/259 J/MNRAS/329/227 J/MNRAS/329/87 J/MNRAS/333/100

J/MNRAS/339/652 J/MNRAS/350/1485 J/MNRAS/352/939 J/MNRAS/369/1351 J/MNRAS/372/425 VII/217 VII/249

Colless M.M.: J/A+AS/133/325 J/MNRAS/292/499 J/MNRAS/321/277 VII/226 VII/250 Collier Cameron A.: J/A+A/467/785 Collier J.: J/ApJ/617/1091 Collier S.: J/ApJS/115/185 Collin B.: J/other/APh/26.282 Collin S.: J/A+A/376/775 J/A+A/422/925 Collinge M.J.: J/AJ/126/1023 J/AJ/126/2125 J/AJ/126/2209 J/AJ/129/2542 J/ApJ/607/426

Collins C.: J/A+A/342/1 J/A+AS/130/323 J/ApJ/514/148 J/MNRAS/329/227 J/MNRAS/329/87

J/MNRAS/333/100 J/MNRAS/350/1485 J/MNRAS/352/939 Collins C.A.: II/211 J/A+A/414/23 J/A+A/425/367 J/ApJS/126/209 J/ApJS/140/239 J/MNRAS/274/1071

J/MNRAS/300/417 J/MNRAS/341/1093 COLLINS C.A.: VII/175 Collins N.R.: J/ApJ/548/127 J/ApJS/146/209 Collister A.: J/MNRAS/375/68 Collmar W.: J/A+AS/143/145 Colom P.: J/A+A/393/1053 Colomb F.R.: VIII/9 Colome J.: J/A+A/390/173 Colomer F.: J/A+A/319/235 Colon C.: J/A+A/422/1109 Colpi M.: J/ApJ/634/272 COLUZZI R.: II/130 Coluzzi R.: III/212 COLUZZI R: VI/71A Coluzzi R.: VI/84 Comastri A. and Mignoli M.: J/A+A/327/61 Comastri A.: J/A+A/370/900 J/A+A/409/79 J/A+A/432/69 J/A+A/457/501 J/A+A/466/31 J/ApJ/570/100

J/MNRAS/327/771 J/MNRAS/342/575 COMBES F.: J/A+A/297/28

Combes F.: J/A+A/382/513 J/A+A/386/140 J/A+A/412/657 J/A+A/438/507 J/A+A/465/1099

J/A+A/470/505 J/A+A/472/121 J/AJ/121/225 Combi J.A.: J/A+A/370/468 Comeron F.: J/A+A/375/539 J/A+A/425/489 J/A+A/473/149 J/A+AS/146/323 COMINSKY L.: VII/18 Comito C.: J/A+A/469/207 J/ApJS/156/127 Commons K.: J/ApJS/126/209 Comoretto G.: J/A+A/362/199 J/A+A/368/845 J/A+AS/101/153 J/A+AS/103/541 COMTE G.: J/A+A/285/1 Comte G.: J/A+AS/133/171 Concannon K.D.: J/AJ/131/527 Conconi P.: J/A+A/469/503 Conde H.: III/18B Condon J.J.: J/AJ/102/1663 J/AJ/109/2318 J/AJ/125/2411 J/AJ/127/1959 J/AJ/87/1429 J/AJ/87/739

J/AJ/90/1437 J/AJ/99/1071 J/ApJ/252/102 J/ApJ/378/65 J/ApJ/554/803 J/ApJS/117/361 J/ApJS/123/219 J/ApJS/123/41 J/ApJS/125/409 J/ApJS/129/547 J/ApJS/159/282 J/ApJS/36/53 J/ApJS/65/485 J/ApJS/73/359 J/ApJS/90/173 J/PASP/113/10

CONDON J.J.: VII/97 Condon J.J.: VIII/14 CONDON J.J.: VIII/26 Condon J.J.: VIII/34 VIII/40 VIII/6 VIII/61 VIII/65 VIII/79 Condon M.A.: J/AJ/87/739 J/ApJ/252/102 Confino B.: J/A+A/356/590 Conidis G.: J/AJ/130/767 Conley A.: J/A+A/437/789 Connaughton V.: IX/20A J/ApJS/154/585 Connelley M.S.: J/ApJ/602/571 Conner S.: J/ApJS/75/801 J/AZh/73/835 VIII/52B Conner S.R.: VIII/52B Connolly A.: J/AJ/123/2223 J/ApJS/159/41 J/PASJ/55/771 Connolly A.J.: J/AJ/121/2308 J/AJ/121/2331 J/AJ/123/567 J/AJ/125/1817 J/AJ/130/968 J/AJ/131/84 Connors A.: J/A+AS/143/145 Conrad A.: J/AJ/127/3121 Conrad J.: J/other/APh/26.282 Conrow T.: II/156A Conroy C.: J/ApJ/625/6 Conroy M.A.: J/AJ/129/363 Conselice C.: J/A+A/389/439 J/MNRAS/358/149 Conselice C.J.: J/A+A/451/1159 J/A+A/459/679 J/AJ/117/75 J/AJ/122/3046 J/AJ/126/1183 J/AJ/127/131

J/AJ/129/682 J/ApJ/603/503 J/ApJS/147/1 CONSIDERE S.: I/74 Considere S.: J/A+AS/110/19 J/A+AS/130/285 CONTARINI G.: J/A+A/283/911 Conti P.S.: II/82 III/85 J/AJ/117/1708 J/ApJ/559/864 J/ApJS/107/281 Contini M.: J/A+A/323/71

Contini T.: J/A+A/416/515 J/A+A/428/1043 J/A+A/448/893 J/A+A/457/79 J/A+A/463/519 J/A+A/468/937 J/A+A/474/473 J/A+AS/130/285 J/A+AS/136/35 J/MNRAS/330/75 J/MNRAS/362/1143 Contreras J.L.: J/A+A/390/39 J/A+A/406/L9 J/A+A/410/813 Contreras Martinez M.E.: J/A+AS/97/951 Cook H.K.: J/A+A/423/97

Cook K.H.: J/AJ/114/326 J/AJ/124/2039 J/AJ/127/334 J/ApJ/463/26 J/ApJ/542/257 J/ApJ/552/289

J/ApJ/562/337 J/ApJ/598/597 Cooke B.: J/MNRAS/274/1194 Cooke B.A.: VII/19 COOKE D.J.: J/MNRAS/251/330 Cooke J.: J/AJ/123/2206 J/MNRAS/358/795 Cooke J.A.: J/A+A/435/363 J/MNRAS/318/1206 Cool A.M.: J/ApJ/632/266 Cool R.: J/AJ/132/231 Cool R.J.: J/ApJ/644/100 Cooley J.: J/other/APh/26.282 Cooper D.: J/MNRAS/334/20 Cooper M.C.: J/AJ/127/3121 J/AJ/131/332 J/ApJ/625/6 J/ApJ/633/174 J/MNRAS/371/221 Copenhagen University Obs.: I/147 I/170A I/189 I/205 I/213 I/256 I/304 Copenhagen University Obs. Royal Greenwich Obs. Instituto y Observatorio de Marina (1984): I/117 Copet E.: J/A+A/349/L69 J/A+A/357/219 J/A+A/372/173 Copetti M.V.F.: J/A+AS/91/285 Copin Y.: J/PASP/109/883 Coppi P.: J/AJ/127/1158 J/MNRAS/359/930 Coppi P.S.: J/AJ/134/1103 J/ApJ/621/104 J/ApJS/162/1 Coppin K.: J/MNRAS/372/1621 Cora A.: J/A+AS/121/119 Corbally C.: J/A+A/345/521 J/AJ/124/2955 Corbally C.J.: III/117 J/AJ/126/2048 J/AJ/132/161 J/ApJ/541/841 Corbard T.: J/A+A/390/717 Corbelli E.: J/other/PASA/19.455 CORBEN J.B.: VIII/45 CORBEN P.M.: I/129 Corbin T.: I/259 CORBIN T.E.: I/160 Corbin T.E.: I/171 I/172 I/219 I/222 I/223 I/224 I/225 I/226 I/227 I/228 I/232 I/246 I/247 I/275 CORBIN T.E.: I/72 Corbin T.E.: J/AJ/110/880 J/AJ/120/1120 Corburn W.: J/A+A/395/129 Corcoran M.: J/ApJS/96/303 Corcoran M.F.: J/A+A/395/499 J/AJ/129/2018 J/ApJ/562/446 J/ApJ/607/904 J/ApJ/616/1284 Cordero-Gracia M.: J/AJ/105/427 Cordes J. M.: J/ApJS/121/171 Cordes J.M.: J/ApJ/605/759 J/ApJS/118/201 Cordes O.: J/A+A/441/905 Cordis L.: J/A+AS/127/145 J/A+AS/128/507 VII/196 Cordoni J.-P.: J/A+AS/46/347 Cordova F.: J/MNRAS/342/802 J/MNRAS/362/1371 J/MNRAS/369/156 Cordova F.A.: J/ApJ/434/54 Corliss C.: VI/74 Corliss Ch.H.: VI/16 Cornell A.P.: J/MNRAS/371/1140 Cornell M.E.: III/166 Cornett R.H.: J/AJ/113/1011 J/AJ/116/180 CORNETT R.H.: J/ApJ/425/122 Cornett R.H.: J/ApJ/438/181 J/ApJS/104/287 CORNETT R.H.: VII/48 Cornide M.: J/A+A/312/221 J/A+A/326/780 J/A+AS/123/473 J/A+AS/146/103 J/MNRAS/328/45 Cornille M.: J/A+AS/105/77 Cornuelle C.S.: II/254 Corongiu A.: J/MNRAS/372/777 Corporon P.: J/A+AS/136/429 Corradi R.: J/A+A/332/721 J/A+A/367/498 Corradi R.L.M.: J/A+A/343/841 J/A+A/379/90 J/A+A/405/627 J/A+A/434/1107 J/A+A/465/481 Corradi W.J.B.: J/A+A/304/415 CORRADI W.J.B.: J/A+AS/112/95 Corradi W.J.B.: J/AJ/126/2971 Corral A.: J/A+A/469/27 J/A+A/474/473 J/A+A/476/1191 CORSI C.E.: J/A+A/290/69 Corsi C.E.: J/A+A/320/757 J/A+A/324/915 J/A+AS/61/473 CORSI C.E.: J/A+AS/67/327 Corsi C.E.: J/AJ/109/650 J/AJ/109/663 J/AJ/118/1671 J/AJ/121/916 CORSI C.E.: J/MNRAS/272/391 Corsico A.: J/A+A/465/249

Corsini E.M.: J/A+A/342/671 J/A+A/360/439 J/A+A/374/394 J/A+A/382/488 J/A+A/408/873

J/A+A/416/507 J/A+A/424/447 J/A+A/440/107 J/A+AS/131/105 Corsini: J/A+A/395/753 Cortes S.: J/PASP/118/1690

Cortese L.: J/A+A/384/383 J/A+A/397/421 J/A+A/401/471 J/A+A/425/429 J/A+A/446/839

J/ApJ/634/272 Cortina J.: J/A+A/390/39 J/A+A/406/L9 J/A+A/410/813 Corwin H.G.: J/PASP/110/779 VII/110A VII/112 VII/116 VII/137B VII/155 VII/4A Corwin H.G.Jr: J/AJ/108/2128 Corwin H.G.Jr.: J/AJ/109/517 Corwin H.G.Jr: J/ApJS/99/391 Corwin H.G.Jr.: J/MNRAS/274/1107 Corwin T.M.: J/AJ/129/267 Cosentino R.: J/A+A/377/123 J/A+A/420/683 J/AJ/121/327 Cossburn M.R.: J/MNRAS/313/347 Costa E.: J/A+A/367/725 J/A+A/455/813 J/A+A/464/573 J/A+A/473/347 J/A+AS/115/75 J/A+AS/139/25

J/A+AS/142/369 J/AJ/112/2408 Costa J.E.S.: J/A+A/468/637 Costa J.M.: J/A+A/363/239 Costa M.: J/PASJ/52/997 Costa R.D.D.: J/A+A/279/567 J/A+A/313/924 Costa R.D.D: J/A+A/380/300 Costa R.D.D.: J/A+A/423/199 J/A+A/427/231 J/A+A/435/1087 J/A+A/472/101 Costa V.: J/A+A/285/404 J/A+A/328/235 J/A+A/407/1059 J/A+A/450/715 J/A+A/456/261 COSTA V.: J/A+AS/106/21 Costa V.: J/A+AS/129/9 J/MNRAS/353/310 Costado M.T.: J/MNRAS/368/253 Costain C.H.: J/A+AS/65/485 J/AJ/74/366 Costero R.: III/84B J/A+A/275/201 J/A+A/349/619 COTE J.: J/A+A/290/609 Cote P.: J/A+A/462/599 COTE P.: J/AJ/106/1508

Cote P.: J/AJ/127/24 J/ApJ/454/788 J/ApJ/559/812 J/ApJ/591/850 J/ApJ/613/279 J/ApJS/153/223

J/ApJS/154/509 J/ApJS/90/83 Cotera A.S.: J/AJ/126/1665 J/AJ/130/1719 J/ApJ/540/1016 Cotter G.: J/MNRAS/354/485 J/MNRAS/379/1343 Cotton W.: J/ApJS/155/421 Cotton W.D.: J/A+A/370/409 J/AJ/125/2411 J/ApJ/591/640 J/ApJS/125/409 VIII/65 VIII/79 Cottrell P.L.: J/MNRAS/362/1167 Couch W.: J/MNRAS/329/227 J/MNRAS/333/100 J/MNRAS/350/1485 J/MNRAS/352/939 J/MNRAS/369/1351

J/MNRAS/370/1223 J/MNRAS/372/425 Couch W.J.: J/A+A/470/39 J/A+A/476/59 J/ApJ/497/188 J/ApJ/542/673 J/ApJ/549/820 J/ApJS/110/213

J/ApJS/122/51 J/MNRAS/283/367 J/MNRAS/287/472 J/MNRAS/327/588 J/MNRAS/329/87 J/MNRAS/335/10 J/MNRAS/366/645 VII/240

Couchot F.: J/A+A/389/149 Coude du Foresto: J/A+A/428/587 Coude du Foresto V.: J/A+A/393/183 J/A+A/433/1155 J/A+A/453/155 Coude Du Foresto V.: J/A+A/459/669 Coudert L.H.: J/A+A/449/855 Coudreau N.: J/A+A/430/373 J/A+A/465/71 Coudreau-Durand N.: J/A+AS/126/471 J/A+AS/130/333 VII/238 Coupinot G.: J/A+A/421/509 Courbin F.: J/A+A/391/481 J/ApJS/135/227 Courteau S.: J/AJ/114/2402 J/ApJS/109/333 VII/198 Courtes G.: J/A+A/268/419 J/A+A/309/446 J/A+AS/31/439 Courtney N.J.D.: J/MNRAS/339/913 Courvoisier T.J.-L.: J/A+A/411/L59 J/A+A/467/585 J/ApJ/635/214 Cousineau S.: J/ApJ/511/595 Cousins A.W.J.: II/222 II/50 J/other/SAAOC/15.30 COUTEAU P.: I/173 Couteau P.: I/209A J/A+AS/106/377 J/A+AS/126/1 Coutinho L.H.: J/ApJS/121/591 Couto da Silva T.C.: J/A+AS/110/371 COUTURE J.: J/ApJS/73/671 Coutures C.: J/A+A/303/137 J/A+A/389/149 J/A+A/421/509 J/A+A/454/185 Covarrubias: J/A+A/445/939 Covarrubias R.: J/AJ/112/2408 J/AJ/130/2453 J/AJ/131/1648 Covey K.R.: J/AJ/123/3409 J/AJ/126/3007 J/AJ/128/426 J/AJ/128/463 J/AJ/130/1145 J/AJ/130/1871

J/AJ/133/439 COVINO E.: J/A+A/299/89

Covino E.: J/A+A/319/184 J/A+A/320/185 J/A+A/353/1044 J/A+A/375/130 J/A+A/433/151

J/A+A/470/281 J/A+AS/114/109 J/A+AS/124/449 J/A+AS/146/323 Covino S.: IX/34 J/A+A/330/423 J/A+A/363/29 J/A+A/367/759 J/A+A/391/179 J/A+A/469/503

J/A+A/474/793 Covone G.: J/A+A/428/339 J/MNRAS/358/1116 J/MNRAS/376/125 Cowan J.J.: J/ApJ/530/783 J/ApJ/573/306 J/ApJ/617/1091 J/ApJ/645/613 J/ApJS/162/227 Cowen D.F.: J/other/APh/26.282 Cowen J.J.: J/AJ/104/704 Cowie L.L.: J/A+A/413/843 J/AJ/122/2810 J/AJ/126/539 J/AJ/126/632 J/AJ/127/180 J/AJ/127/3121

J/AJ/127/3137 J/AJ/128/1483 J/AJ/128/1501 J/ApJ/471/L5 J/ApJ/538/29 J/ApJ/554/742 J/ApJS/94/461 Cowley A.P.: III/31B COWLEY A.P.: J/AJ/104/1706 Cowley A.P.: J/AJ/114/2353 COWLEY A.P.: VII/143 VII/143 VII/143 Cowley C.R.: J/A+A/339/165 Cox A.N.: J/ApJ/119/188 COX C.V.: J/AJ/107/1022 Cox N.: J/A+A/429/559 Cox P.: J/ApJS/160/272 Coyne G.V.: J/A+A/454/277 COYNE G.V.: J/A+AS/102/343 COZIOL R.: J/AJ/105/35 Coziol R.: J/AJ/108/405 J/MNRAS/344/181 Crabbe C.: J/AJ/122/611 Crabtree D.R.: J/AJ/109/2480 J/AJ/112/491 J/AJ/129/729 Cracraft M.: J/AJ/131/461 Craig J.: J/PASP/107/32 Craig N.: IX/35 J/AJ/108/1843 J/PASP/109/998 Cram L.: J/AJ/131/1216 J/MNRAS/305/297 J/MNRAS/306/708 Cram L.E.: J/ApJ/624/135 J/ApJS/155/1 J/MNRAS/352/1245 VIII/16 CRAMER N.: J/A+A/290/609 Cramer N.: J/A+A/325/167 J/A+A/406/373 CRAMPTON D.: III/29 J/AJ/104/1706 Crampton D.: J/AJ/114/2353 J/AJ/127/2455 J/AJ/130/445 J/ApJ/464/79 J/ApJ/481/49 J/ApJ/499/112

J/ApJ/517/148 CRAMPTON D.: VII/143 VII/143 VII/143 Crampton D.: VII/225A Crane J.D.: J/AJ/128/245 Crane P.: J/A+A/276/25 J/A+A/287/769 J/A+A/288/413 Crapsi: J/A+A/475/915 Crause L.: J/A+A/434/1107 Crawford C.S.: J/A+A/476/1191 J/MNRAS/274/75 J/MNRAS/301/881 J/MNRAS/306/857 J/MNRAS/318/333 Crawford D.F.: VIII/2 CRAWFORD D.L.: II/138 Crawford F.: J/AZh/73/835 J/MNRAS/328/17 J/MNRAS/335/275 J/MNRAS/342/1299 J/MNRAS/352/1439

J/MNRAS/372/777 Crawford F.L.: J/A+AS/139/135 Crawford J.: J/MNRAS/308/897 Crawford M.: J/MNRAS/372/1621 Crawford M.D.: J/ApJ/297/582 Creech-Eakman M.: J/AJ/121/2189 Creese M.: J/AJ/110/268 Creevey O.: J/A+A/468/637 Cremades H.: J/A+A/422/337 Cremonese G.: J/A+A/461/789 Crenshaw D: J/A+A/369/57 Crenshaw D.M.: J/AJ/132/321 J/ApJS/150/165 Crespo da Silva L.: J/A+A/465/331 Crew G.: J/other/GCN/1070.1 CREZE M.: I/74 Creze M.: J/A+A/265/32 J/MNRAS/351/1071

Crifo F.: I/196 J/A+A/331/81 J/A+A/346/785 J/A+A/380/590 J/A+A/401/959 J/A+A/411/447

J/A+A/441/653 J/A+A/462/795 Crimier N.: J/A+A/472/1029 CRINKLAW G.: J/PASP/103/536 Crippa R.: J/A+A/465/331 Criscienzo M.D.: J/MNRAS/365/1357 J/MNRAS/371/1503 Crisp D.: J/AJ/113/1001 J/AJ/117/1700 J/PASP/108/682 Cristia S.: J/A+A/373/757 Cristiani S.: J/A+A/318/347 J/A+A/321/123 J/A+A/341/641 J/A+A/375/1 J/A+A/392/395 J/A+A/434/53

J/A+A/437/883 J/A+A/449/951 J/A+A/454/423 J/A+A/478/83 CRISTIANI S.: J/A+AS/112/347

Cristiani S.: J/A+AS/127/217 J/A+AS/140/351 J/AJ/119/2540 J/AJ/120/2206 J/AJ/122/2190

J/AJ/124/2955 J/ApJ/600/L99 J/MNRAS/335/555 Cristobal-Hornillos D.: J/ApJS/154/155 Crochot A.: J/A+A/391/767 Crocker D.A.: J/AJ/118/2071 J/AJ/121/225 Crocker J.H.: J/AJ/121/2308 J/AJ/121/2331 Crockett H.: J/MNRAS/351/1290 Crockett N.: J/PASP/118/1690 Croft S.: J/A+A/461/823 J/MNRAS/373/1531 Croke B.F.: J/A+AS/122/507 Croke B.F.W.: J/MNRAS/341/534 Croom S.: J/MNRAS/372/425 J/MNRAS/375/68 Croom S.M.: II/268 J/ApJ/505/L7 J/MNRAS/276/33 J/MNRAS/306/592 J/MNRAS/328/150 J/MNRAS/328/805

J/MNRAS/335/673 VII/223 VII/241 Cropper M.: J/MNRAS/246/54 Cropper M.S.: J/A+A/476/1191 Crosby N.: J/A+AS/130/233 Crosby T.: J/ApJS/157/175 Cross N.: J/ApJ/621/53 J/MNRAS/329/87 Cross N.J.G.: J/AJ/128/1990 J/AJ/130/1516 J/ApJ/612/202 J/ApJS/150/1 VII/240 VII/240 Crotts A. P. S.: J/ApJS/122/355 Crotts A.P.S.: J/AJ/119/2571 Crovisier J.: J/A+A/393/1053 Crow M.: J/A+A/465/331 CROWNE A.H.: VII/17 VII/33 Crowne A.H.: VIII/3 Crowther P.A.: J/A+A/439/265 Cruddace R.: J/ApJ/514/148 J/ApJS/140/239 Cruddace R.G.: J/A+A/425/367 CRUDDACE R.G.: J/ApJS/56/403 Cruz K.L.: J/AJ/123/2806 J/AJ/123/2822 J/AJ/123/2828 J/AJ/126/2421 J/AJ/126/3007 J/AJ/127/2856

J/AJ/128/463 J/AJ/132/891 J/AJ/133/2258 J/AJ/133/439 Cruz M.J.: J/MNRAS/355/20 J/MNRAS/373/1531 J/MNRAS/375/1349 Cruzalebes P.: J/A+A/465/469 Cruz-Gonzales C.: III/84B Cruz-gonzalez I.: J/A+A/349/619 Cruz-Gonzalez I.: J/A+A/361/660 J/A+A/399/1135 J/A+AS/104/271

Csabai I.: J/AJ/121/2308 J/AJ/121/2331 J/AJ/121/31 J/AJ/122/503 J/AJ/123/567 J/AJ/124/1651

J/AJ/125/1817 J/ApJ/598/260 J/MNRAS/371/495 Csak B.: J/A+A/374/712 J/A+A/375/285 J/A+A/376/561 J/A+A/394/97 J/A+A/463/589 J/other/P+SS/53.925 Cseresnjes P.: J/A+A/357/871 J/A+A/375/909 Csizmadia Sz.: J/A+A/403/637 J/A+A/417/745 J/A+A/426/1001 J/A+A/465/943 J/AN/325/424 Cuadra J.: J/ApJ/643/1011 Cubaynes D.: J/A+A/439/387 CUCCHIARO A.: III/53 Cucciati O.: J/A+A/430/411 J/A+A/457/79 Cudworth K.M.: J/AJ/105/1822 J/AJ/110/2747 J/AJ/121/935 J/ApJ/420/612 J/ApJ/443/124 J/ApJ/549/578 Cudworth Kyle M.: J/ApJ/347/251 CUESTA L.: J/A+AS/104/169 Cuevas H.: J/A+A/377/428 J/A+A/407/31 CUEVAS H.: J/A+AS/104/1 Cuevas H.: J/ApJS/126/1 Cuffey J.: J/MNRAS/122/363 Cuillandre J.-C.: J/A+A/372/391 J/A+A/400/891 J/A+A/422/407 Cuillandre J.C.: J/A+A/448/189 Cuillandre J.-C.: J/A+A/450/147 J/A+A/451/1159 Cuillandre J.C.: J/A+A/471/499 Cuillandre J.-C.: J/ApJ/610/1064 Cuisinier F.: J/A+A/307/215 J/A+A/423/133 J/A+AS/125/229 J/A+AS/130/65 Cullen M.F.: II/37 Cunba K.: J/A+A/448/1001 Cunha K.: J/A+A/420/183 J/A+A/441/1149 J/ApJS/151/313 Cunha N.C.S.: J/A+AS/128/251 Cunningham M.: J/MNRAS/367/1609 Cunningham M.R.: J/MNRAS/368/1223 CUNOW b.: II/200 CUNOW B.: J/A+AS/107/277 Cunow B.: J/A+AS/125/71 J/A+AS/129/593 CURCHOD A.: III/56 Curdt W.: J/A+A/375/591 J/A+AS/126/281 Curiel S.: J/ApJS/125/427 Curiel Z.: J/A+A/387/1146 Curioni G.P.: J/A+AS/103/541 Curioni P.: J/A+AS/103/541 Curran J.: VIII/70A VIII/78 VIII/81 Curran: J/A+A/385/1014 Curran S.J.: J/other/PASA/19.455 Currie M.J.: J/A+AS/127/367 J/MNRAS/279/595 J/MNRAS/373/781 Curry S.: J/ApJ/610/247 Custiani S.: VII/69 Custo A.: J/AJ/129/1534 Cusumano G.: J/A+A/402/647 J/A+A/448/L9 J/A+A/474/793

Cutispoto G.: J/A+A/364/205 J/A+A/367/910 J/A+A/384/491 J/A+A/400/659 J/A+A/432/657 J/A+A/469/503 J/A+AS/115/41 J/A+AS/119/281 J/A+AS/121/369 J/A+AS/125/11 J/A+AS/127/207 J/A+AS/131/197 J/A+AS/131/321 J/A+AS/138/87

Cutri R.: J/AJ/125/2521 J/ApJ/560/566 J/ApJ/635/214 J/ApJS/95/1 J/PASP/113/10 Cutri R.M.: II/126 II/246 J/AJ/112/62 J/AJ/126/63 J/AJ/127/646 J/ApJ/564/421 J/ApJ/569/23

Cuypers J.: J/A+A/304/415 J/A+A/402/1157 J/A+A/437/705 J/A+A/438/201 J/A+A/456/599

J/A+A/463/243 J/A+A/475/1159 J/A+AS/112/291 J/A+AS/139/425 Cvetkovic Z.: J/A+A/400/791 J/A+A/425/1147 J/other/BOBeo/155.3 Czoske O.: J/A+A/372/391 J/A+A/441/905 J/ApJ/591/53 J/MNRAS/330/1

D.: J/A+A/395/753 J/A+A/395/753 J/A+A/395/753 J/A+A/428/587 J/A+A/428/587 Da Costa G.S.: J/A+A/448/983 DA COSTA G.S.: J/AJ/106/986 Da Costa G.S.: J/AJ/110/652 J/AJ/119/1748 da Costa G.S.: J/AJ/124/1464 Da Costa G.S.: J/AJ/127/318 J/AJ/129/2232 J/AJ/130/1065 J/AJ/131/912 J/ApJ/297/582 J/ApJ/298/544

J/ApJS/90/83 J/ApJS/99/637 Da Costa L.: J/A+A/379/436 da Costa L.: J/A+A/379/740 J/A+A/384/81 J/A+A/392/741 J/A+A/409/439 J/A+A/435/781 Da Costa L.: J/A+A/442/841 da Costa L.: J/A+A/452/119 J/A+A/456/881 Da Costa L.: J/A+A/462/553 J/A+AS/137/75 J/A+AS/137/83 da Costa L.N.: J/A+A/408/905 J/AJ/108/1987 Da Costa L.N.: J/AJ/113/22 J/AJ/116/1 J/AJ/117/1668 J/AJ/117/2039 da Costa L.N.: J/AJ/123/2976 J/AJ/125/2307 Da Costa L.N.: J/AJ/126/2268 D.A.: J/A+A/364/391 da Silva J.R.P.: J/ApJ/578/943 da Silva L.: J/A+A/434/275 J/A+A/440/1153 J/A+A/441/1149 J/A+A/458/895 J/A+A/460/695 Da Silva L.: J/AJ/126/2971 Da Silva Neto D.N.: J/AJ/125/2728 da Souza R.: J/A+A/258/243 Dachs J.: J/A+A/312/818 Dacic M.: J/A+A/400/791 J/A+A/425/1147 J/other/BOBeo/155.3 D’Addario L.: J/AJ/119/2801 Daddi E.: J/A+A/392/395 J/A+A/432/69 J/A+A/437/883 J/A+A/465/1099 J/AJ/125/1107 J/ApJ/604/521 Dadina M.: J/ApJS/121/473 Daflon S.: J/A+A/448/1001 Dagkesamanskii R.D.: J/AZh/77/21 D’agostino M.V.: J/other/APh/26.282 Dahlem M.: J/A+A/473/399 J/A+AS/127/145 Dahlen T.: J/ApJ/613/200 Dahlgren M.: J/A+AS/113/115 Dahlmark L.; Skiff B.A.: II/238 Dahm S.E.: J/AJ/123/304 J/AJ/128/1233 J/AJ/129/829 J/AJ/130/1805 J/AJ/131/1530 Dahn C.C.: J/AJ/103/638 J/AJ/105/1571 J/AJ/126/1023 J/ApJ/564/421 J/ApJ/607/426 Dai X.: J/ApJ/625/633 Daigle O.: J/A+A/466/905 Daigne G.: I/272 I/303 J/A+A/437/1135 Dalcanton J.: J/A+A/444/365 Dalcanton J.J.: J/A+A/427/397 J/ApJ/563/629 J/ApJ/566/103 J/ApJS/137/117 J/MNRAS/351/265 Dale D.A.: J/AJ/118/1468 J/AJ/118/1489 J/ApJ/633/871 D’Alessio F.: J/A+A/469/503 d’Alessio F.: J/PAZh/30/243 D’Alessio P.: J/ApJ/638/897

Dallacasa D.: J/A+A/363/887 J/A+A/369/380 J/A+A/379/755 J/A+A/419/71 J/A+A/426/463

J/A+A/450/959 J/A+A/463/519 J/A+AS/131/303 Dallaporta S.: J/A+A/452/567 Dallier R.: J/A+AS/116/239 Dall’Ora M.: J/A+A/428/339 Dallora M.: J/A+A/462/599 Dalton G.: J/MNRAS/329/227 J/MNRAS/333/100 J/MNRAS/350/1485 J/MNRAS/352/939 Dalton G.B.: J/AJ/123/51 J/MNRAS/269/151 J/MNRAS/269/301 J/MNRAS/289/263 J/MNRAS/329/87 Daly R.A.: J/ApJ/612/652 Dambis A.K.: J/A+A/323/775 J/A+AS/143/211 J/PAZh/25/689 J/PAZh/27/132 D’Ambrosio V.: J/A+AS/125/399 Dame T.M.: J/A+A/357/1001 VIII/39 Damerdji Y.: J/AJ/133/1470

Damiani F.: J/A+A/341/751 J/A+A/355/651 J/A+A/421/175 J/A+A/430/941 J/A+A/450/993 J/A+A/460/133 J/A+A/462/123 J/A+A/462/245 J/A+A/464/211 J/ApJ/582/382 J/ApJ/588/1009 J/ApJ/606/466 J/ApJ/608/781 J/ApJS/160/319

D’Amico N.: J/MNRAS/328/17 J/MNRAS/335/275 J/MNRAS/342/1299 D’amico N.: J/MNRAS/352/1439 D’Amico N.: J/MNRAS/372/777 Damineli A.: J/A+A/465/993 J/A+AS/133/299 Dan M.: B/iram

Danese L.: J/A+A/463/455 J/AJ/127/3075 J/MNRAS/302/222 J/MNRAS/305/297 J/MNRAS/325/1173

J/MNRAS/341/993 J/MNRAS/351/1290 Daniel K.J.: J/ApJ/613/393 Daniel S.A.: J/PASP/106/281 Daniels J.: J/MNRAS/324/580 Danks A.C.: III/179 Danner R.: J/A+A/310/384 J/A+AS/128/331 Dannerbauer H.: J/A+A/465/1099 J/ApJ/606/664 Dantas C.C.: J/ApJ/485/447 Dantel M.: J/A+A/444/365 Dantel-Fort M.: J/A+A/422/407 D’Antona F.: J/A+A/464/939 J/A+A/464/953 Danziger I.J.: J/A+AS/120/257 J/MNRAS/363/769 J/other/SerAJ/166.13Short Danziger J.: J/A+A/469/503 J/AJ/112/2408 DANZINGER I.J.: III/22 Daou D.: I/176

Dapergolas A.: J/A+A/369/932 J/A+A/374/164 J/A+A/447/89 J/A+AS/116/1 J/A+AS/136/525

J/MNRAS/351/1290 Dappen W.: J/A+AS/135/405 Darling J.: J/AJ/124/100 Dasgupta A.K.: J/A+A/388/149 J/A+A/458/767 Dashevsky I.: J/AJ/119/486 Daszynska-Daskiewicz J.: J/A+A/463/243 Daszynska-Daszkiewicz J.: J/A+A/415/241 J/MNRAS/353/310 Daugherty D.A.: J/ApJS/95/457 Dautov I.: J/A+A/375/670 D’Avanzo P.: J/A+A/474/793 Dave Monet: I/243 I/252 Dave R.: J/ApJ/636/631 Davenhall A.C.: VI/49 Davey A.R.: J/A+AS/133/257 J/A+AS/142/85 David L.P.: J/ApJ/412/479 J/ApJ/519/533 J/ApJ/635/1053 David P.: J/A+A/277/453 J/A+A/293/833 J/A+AS/111/237 Davidge T.J.: J/AJ/112/2559 J/AJ/115/130 J/AJ/132/271 Davidsen A.F.: J/AJ/124/1651 J/ApJ/475/469 Davidson J.W.JR: J/AJ/131/1000 Davies B.: J/A+A/439/1107 Davies J.: J/A+A/350/883 J/A+A/372/173 J/A+A/420/945 J/A+A/421/623 J/A+A/430/411 J/MNRAS/352/478

J/MNRAS/357/819 Davies J.G.: VII/8 Davies J.I.: J/A+A/357/219 J/MNRAS/304/297

Davies J.K.: J/A+A/365/110 J/A+A/378/116 J/A+A/379/564 J/A+A/412/405 J/AJ/112/2274

J/MNRAS/373/781 Davies M.B.: J/A+A/459/489 Davies R. D.: J/A+A/397/213 Davies R.: J/AJ/130/445 Davies R. L.: J/MNRAS/326/473 Davies R.D.: J/A+A/370/78 J/MNRAS/341/1066 J/MNRAS/360/340 VII/232 Davies R.I.: J/A+A/406/593 DAVIES R.L.: J/AJ/100/1091

Davies R.L.: J/AJ/128/1558 J/ApJS/106/1 J/ApJS/143/73 J/MNRAS/292/499 J/MNRAS/305/259 J/MNRAS/313/469 J/MNRAS/321/277 J/MNRAS/327/249 J/MNRAS/327/265 J/MNRAS/337/256 J/MNRAS/358/233 VII/176A

Davignon G.: J/A+A/449/305 J/A+A/463/243 DAVILA H.: J/A+AS/103/427 Davila J.M.: J/ApJ/543/1016 Davis C.J.: J/MNRAS/353/1025 J/MNRAS/359/2 J/MNRAS/373/781 Davis D.S.: J/AJ/123/1216 Davis J.: J/A+A/455/259 Davis: J/A+A/385/1014 DAVIS J.R.: II/6 DAVIS L.E.: J/AJ/100/1091 DAVIS M.: II/174 Davis M.: J/AJ/121/31 J/AJ/127/3121 J/ApJ/475/445 J/ApJ/625/6 J/ApJ/633/174 J/MNRAS/371/221

VII/185 VII/185 VII/185 Davis M.M.: J/AJ/77/405 Davis R.J.: J/AJ/124/1144 J/MNRAS/341/1066 J/MNRAS/360/340 J/PASP/118/1656 Davis S.: J/AJ/127/2771 Davis T.M.: J/AJ/131/1648 Davour A.: J/other/APh/26.282

Davoust E.: J/A+A/323/337 J/A+A/416/515 J/A+AS/110/19 J/A+AS/116/429 J/A+AS/126/67

J/A+AS/130/285 J/A+AS/141/113 DAVOUST E.: VII/39A Dawson B. R.: J/ApJS/121/171 Dawson D.W.: J/A+A/374/1056 J/AJ/115/1076 Dawson P.C.: J/PASP/117/1 Dawson S.: J/AJ/122/598 J/AJ/123/2223 J/AJ/130/1324 J/ApJ/600/L99 Day C.T.: J/other/APh/26.282 Dayal A.: J/ApJS/122/221 De Almeida M.L.T.: J/AJ/130/1111 de Angeli F.: J/A+A/391/945 De Angeli F.: J/A+A/432/851 J/A+A/440/901 J/A+A/459/489 de Angelis G.: J/A+AS/113/115 De Angelis G.: J/A+AS/125/399 de Batz B.: J/A+A/363/62 De Becker M.: J/A+A/463/981 De Benedetto G.: J/A+AS/125/399 de Blok W.J.G.: J/A+A/385/816 J/AJ/123/873 J/AJ/124/1954 J/AJ/124/690 De Blok W.J.G.: J/AJ/128/16 J/ApJ/613/914 de Blok W.J.G.: J/MNRAS/337/641 De Blok W.J.G.: J/MNRAS/346/787 J/MNRAS/355/1303 De Boer K.S.: J/A+A/328/167 J/A+A/442/229 DE BOER K.S.: J/A+AS/112/367 de Boer K.S.: J/MNRAS/327/23 de Boer K.S. Seitter W.C.: I/221

De Breuck C.: J/A+A/362/519 J/A+A/431/793 J/A+A/461/823 J/A+AS/143/303 J/AJ/123/637

J/ApJS/131/185 J/MNRAS/347/837 de Bruijne J.H.J.: J/A+A/367/111 De Bruijne J.H.J.: J/AJ/117/354 de Bruijne J.H.J.: J/MNRAS/310/585 de Bruyn A.G.: J/A+A/374/861 J/A+AS/124/259 J/ApJS/149/123 VIII/72 de Carvalho R. R.: J/AJ/119/1562 De Carvalho R.R.: J/A+A/445/857 de Carvalho R.R.: J/A+AS/110/371 De Carvalho R.R.: J/AJ/116/1573 de Carvalho R.R.: J/AJ/125/1660 J/AJ/125/2064 De Carvalho R.R.: J/AJ/128/1017 J/AJ/130/425 de Carvalho R.R.: J/ApJ/485/447 J/ApJS/110/1 De Castro E.: J/A+A/312/221 J/A+A/326/780 J/A+AS/123/473 J/A+AS/146/103 J/MNRAS/328/45 de Cat P.: J/A+A/393/965 De Cat P.: J/A+A/438/201 J/A+A/446/583 J/A+A/449/305 J/A+A/463/243 J/A+A/463/579 De Clercq C.: J/other/APh/26.282 De Cuyper J.-P.: J/A+A/331/949 J/A+AS/128/409 de Franca J.A.Jr: J/A+A/313/924 De Francesco G.: J/A+A/327/61 J/A+AS/121/119 J/A+AS/127/445 J/A+AS/129/577 J/A+AS/130/109

J/A+AS/132/361 J/A+AS/144/481 de Freitas Pacheco J.A.: J/A+A/279/567 J/A+A/313/924 J/A+A/383/30 de Geus E.: VIII/39 DE GEUS E.J.: J/A+A/289/101 J/A+AS/85/915 De Geus E.J.: J/ApJ/428/693 De Gioia-Eastwood K.: J/ApJ/454/151 de Graauw T.: J/A+A/276/25 J/A+A/390/1033 De Grandi S.: J/A+A/425/367 J/A+AS/140/155 J/ApJ/514/148 de Grandi S.: J/ApJS/140/239 De Gregorio-Monsalvo I.: J/ApJ/628/789 DE GRIJP M.H.K.: J/A+AS/108/61 de Grijp M.H.K.: J/A+AS/70/95 J/A+AS/96/389 De Grijp M.H.K.: J/ApJS/158/139 de Grijs R.: J/AJ/121/768 de GROOT M.: II/200 de Groot M.: J/A+A/304/415 De Groot M.: J/A+A/343/847 De Heij V.: J/A+A/391/159 de Heij V.: J/AJ/123/873 DE JAGER C.: III/141 De Jong R.S.: J/A+A/313/377 de Jong R.S.: J/A+A/414/23 DE JONG R.S.: J/A+AS/106/451 De Jong R.S.: J/A+AS/118/557 J/ApJS/162/80 de Jong T.: J/A+A/319/235 J/A+AS/101/513 DE JONG T.: J/A+AS/104/179 DE JONGE A.R.W.: II/148 de Jonge A.R.W.: II/97 DE JUAN L.: J/ApJS/91/507 de Kat J.: J/A+A/389/149 De Kat J.: J/A+A/454/185 de Koff S.: II/202 J/A+A/304/415 J/A+AS/128/251 J/ApJS/122/81 de Koter A.: J/A+A/384/473 De Koter A.: J/A+A/437/467 J/A+AS/106/141 de la Cruz C.: J/A+AS/129/9 de la Fuente Marcos C.: J/A+A/466/931 de la Fuente Marcos R.: J/A+A/428/67 J/A+A/466/931 de la Raze R.: J/A+A/448/1001 de la Reza R.: J/A+A/460/695 De La Torre S.: J/A+A/457/79 de Lapparent V.: J/A+A/475/791 de Lara E.: J/A+AS/145/187 J/other/RMxAA/38.141 de Laverny P.: J/A+A/423/353 J/A+A/466/L1 J/A+AS/122/415 de Lee N.: J/AJ/125/1383 de Lignie M.: J/A+AS/128/179 de Loore C.: J/A+AS/103/67 de los Reyes R.: J/A+A/410/813 De Lucia F.C.: J/ApJ/499/517 J/ApJ/500/1059 J/ApJS/158/188 J/ApJS/162/428 De Lucia G.: J/A+A/427/397 J/A+A/428/339 J/A+A/444/365 De Marchi F.: J/A+A/459/489 J/A+A/471/515 de Marchi G.: J/AJ/121/3114 DE MARCHI G.: J/ApJS/85/293 De Marco O.: J/AJ/123/269 De Martino D.: J/MNRAS/373/687 De Medeiros J.R.: J/A+A/361/1011 J/A+A/363/239 de Medeiros J.R.: J/A+A/364/205 J/A+A/384/491 J/A+A/384/491 J/A+A/395/97 De Medeiros J.R.: J/A+A/427/313 de Medeiros J.R.: J/A+A/458/895 De Medeiros J.R.: J/A+A/466/L1 de Medeiros J.R.: J/A+AS/139/433 J/ApJ/578/943 De Meester W.: J/A+A/463/243 De Mello D.: J/ApJ/613/200 de Mello D.F.: J/A+AS/125/329 De Mello D.F.: J/AJ/118/1131 J/AJ/120/2747 DE MELLO D.F.: VII/205 De Mey K.: J/A+A/310/164 De Oliveira C.M.: J/AJ/121/148 J/ApJ/518/262 J/ApJ/621/673 de Oliveira W.M.: J/A+AS/117/335 de Oliveira-Abans M.: J/A+AS/129/357 De Pasquale M.: J/A+A/455/813 DE PAUW M.: J/A+A/286/136 De Pree C.G.: J/ApJS/155/577 De Propris R.: J/AJ/130/1516 de Propris R.: J/ApJS/142/153 J/MNRAS/329/227 J/MNRAS/329/87 J/MNRAS/333/100 De Propris R.: J/MNRAS/350/1485 J/MNRAS/352/939 J/MNRAS/372/425 J/MNRAS/375/68 VII/240 De Ridder J.: J/A+A/449/305 J/A+A/463/243 J/A+A/465/965 De Robertis M.M.: J/AJ/131/1236 J/PASP/117/1 de Ruiter H.: J/A+A/403/857 de Ruiter H.R.: J/A+A/339/34 De Ruiter H.R.: J/A+A/439/487 J/A+A/457/517 de Ruiter H.R.: J/A+AS/106/1 J/A+AS/95/1 VIII/63 De Ruyter S.: J/A+A/463/243 De Sanctis M.C.: J/A+A/461/789 De Santis C.: J/A+A/449/951 DE SANTIS R.: J/A+AS/101/195 de Sierra A.C.: III/18B de Souza R.: J/A+AS/101/475 DE SOUZA R.: J/A+AS/110/279 de Souza R.E.: J/A+AS/110/371 J/A+AS/125/329 De Ugarte Postigo A.: J/ApJ/641/L13 de Vaucouleurs A.: J/AJ/109/517 VII/112 VII/116 VII/137B VII/155 VII/16 DE VAUCOULEURS A.: VII/167 de Vaucouleurs G.: J/AJ/108/2128 J/AJ/109/517 De Vaucouleurs G.: J/MNRAS/274/1107 de Vaucouleurs G.: VII/112 VII/116 VII/137B VII/155 VII/16 DE VEGT C.: I/151 I/169 De Vegt C.: I/265 de Vegt C.: I/273A J/A+AS/122/529 J/AJ/110/880 V/67 de Vegt Ch.: I/85 de Vicente P.: J/A+A/455/971 De Vries W.H.: J/A+A/463/97 de Vries W.H.: J/AJ/123/1784 J/AJ/127/2565 J/AJ/131/666 de Winter D.: J/A+A/365/110 J/A+A/378/116 J/A+A/379/564 J/A+A/380/609 J/A+AS/104/315 DE WINTER D.: J/A+AS/105/311 De Winter D.: J/A+AS/119/1 de Winter D.: J/A+AS/121/223 J/A+AS/123/63 J/A+AS/124/33 de Wit W.J.: J/A+A/448/189 de Wit W.-J.: J/A+A/450/147 DE YONG T.: VII/109 DE ZEEUW P.T.: J/A+A/289/101 de Zeeuw P.T.: J/A+A/360/472 J/A+A/367/111 J/A+A/445/503 J/A+AS/120/257 De Zeeuw P.T.: J/AJ/117/354

De Zotti G.: J/A+A/415/549 J/A+A/445/465 J/A+A/454/453 J/A+AS/141/257 J/MNRAS/354/305

J/MNRAS/371/898 DE ZOTTI G.: VII/181 Deacon N.R.: J/A+A/416/125 J/A+A/435/363 J/A+A/468/163 J/MNRAS/380/712 Deacon R.M.: J/ApJS/155/595 Dean A.J.: J/ApJ/607/L33 J/ApJ/636/765 J/ApJS/170/175 Dearborn D.S.P.: J/AJ/121/362 Deb N.C.: J/ApJS/121/265 Debehogne H.: J/A+A/304/415 J/A+AS/113/115 Debernardi Y.: J/A+A/354/881 J/A+A/361/991 J/A+A/374/204 J/A+A/470/1009 Debond H.: J/AJ/130/767 Debosscher J.: J/A+A/463/243 J/A+A/475/1159 Decin G.: J/A+A/463/243 Decin L.: J/A+A/390/1033 J/A+A/463/243 J/A+A/463/579 Dedecker M.: J/A+A/379/305 Deeg H.J.: J/A+A/365/110 J/A+A/378/116 J/A+A/379/564 J/A+AS/119/221 J/A+AS/129/455 Deeley K.: J/MNRAS/329/87 Degioia-Eastwood K.: J/AJ/105/1822 DeGioia-Eastwood K.: J/AJ/119/2214 Degioia-Eastwood K.: J/ApJ/549/578 Degirmenci O.L.: J/A+A/409/959 Degli Esposti C.: J/A+A/425/767 J/A+A/455/1161 Degl’Innocenti S.: J/A+A/462/107 VI/113 Deguchi S.: J/A+A/470/957 J/ApJS/106/463 J/PASJ/53/293 J/PASJ/53/517 J/PASJ/54/719 J/PASJ/55/203

J/PASJ/55/229 J/PASJ/56/261 J/PASJ/58/529 Degychi S.: J/PASJ/56/765 Deharveng J.: J/ApJ/504/113 Deharveng J.M.: J/A+A/287/769 Deharveng J.-M.: J/A+A/288/413

Deharveng L.: J/A+A/344/943 J/A+A/347/841 J/A+A/360/1107 J/A+A/381/941 J/A+A/394/225

J/A+A/399/1135 J/A+A/422/129 Deines-Jones P.: J/ApJS/123/79 DEJAIFFE R.: I/47 Dejonghe H.: J/A+A/430/165 J/A+AS/120/257 Dejonghe H.B.: J/ApJ/414/454 Dekany I.: J/A+A/430/1049 Dekel A.: J/ApJS/109/333 VII/198 Del Burgo C.: J/MNRAS/351/1290 del Olmo A.: J/A+A/285/404 del Peloso E.F.: J/A+A/434/275 J/A+A/440/1153 J/A+A/441/1149 Del Principe M.: J/AJ/129/2714 del Rio G.: J/A+A/307/80 del Rizzo D.: J/PASJ/52/997 Del Santo M.: J/ApJ/607/L33

Del Zanna G.: J/A+A/422/731 J/A+A/430/331 J/A+A/432/1137 J/A+A/433/731 J/A+A/459/307

J/A+A/466/763 Delahaye F.: J/A+AS/144/141 Delain K.M.: J/ApJ/599/1067 Delaney M.: J/A+A/407/791 Delaney T.: J/ApJ/613/343 DELEUILE M.: J/A+A/299/557

Delfosse X.: J/A+A/331/581 J/A+A/338/1066 J/A+A/364/665 J/A+A/380/590 J/A+A/401/959

J/A+A/441/653 J/A+A/474/293 J/A+A/475/1125 J/AJ/126/1526 J/ApJ/633/630 Delgado A.J.: J/A+A/421/623 J/A+A/467/1397 DELGADO A.J.: J/AJ/106/1547 Delgado A.J.: J/AJ/116/1801 D’Elia V.: J/A+A/457/501 J/A+A/466/31 Deliyannis C.P.: J/AJ/122/1486 Della Ceca R.: J/A+A/393/425 J/A+A/428/383 J/A+A/469/27 J/A+A/476/1191 J/ApJS/121/473 DELLA CECA R.: VII/181 DELLA PRUGNA F.: J/A+AS/111/229 Della Prugna F.: J/AJ/127/1158 Della Valle A.: J/A+A/454/453 Della Valle M.: J/A+A/338/1006 J/A+A/389/439 J/A+A/459/875 J/AJ/112/2408 Dell’Antonio I.: J/AJ/124/601 Delmas F.: J/A+AS/135/503 J/A+AS/137/273 J/A+AS/137/451 Delorme P.: J/ApJ/633/630 delosReyes R.: J/A+A/406/L9 Delsanti A.C.: J/A+A/380/347 J/A+A/389/641 Demaison J.: J/A+A/466/255 Demarco R.: J/A+A/432/381 J/AJ/131/143 J/ApJ/644/30 Demarque P.: J/AJ/110/652 J/AJ/113/1045 J/AJ/122/1486 DEMARQUE P.: VI/40 Dement’eva A.A.: J/A+AS/118/105

Demers S.: II/4 J/A+A/410/795 J/A+A/416/111 J/A+A/417/479 J/A+A/418/33 J/A+A/424/125

J/A+A/436/91 J/A+A/447/473 J/A+A/456/905 J/A+A/473/143 J/A+AS/122/463 DEMERS S.: J/A+AS/91/171 J/AJ/105/35 Demers S.: J/AJ/108/405 J/AJ/119/2780 J/AJ/120/1801 J/AJ/123/832 J/AJ/125/1298 J/AJ/125/3037 DEMERS S.: J/MNRAS/261/657 DEMICRAN O.: J/A+AS/106/373

Demircan O.: J/A+A/376/158 J/A+A/387/240 J/A+A/403/675 J/A+A/417/263 J/A+A/463/233 J/MNRAS/349/1069 J/MNRAS/357/497 J/MNRAS/370/2013 J/MNRAS/373/1483 J/MNRAS/378/179

Demory B.-O.: J/A+A/471/L51 J/A+A/475/1125 Dempsey J.T.: J/MNRAS/367/1609 Dempsey R.C.: J/A+A/310/173 Demyk K.: J/A+A/466/255 Den Hartog E.A.: J/ApJS/162/227 den Hartog R.: J/A+A/310/31 J/A+A/310/8 J/A+AS/129/399 den Hartog R.H.: J/A+A/229/362 Den Herder J.W.A.: J/A+A/420/135 Denby M.: J/MNRAS/274/1165 Deng J.S.: J/AJ/130/1640 Deng L.: J/A+A/455/247 J/AJ/117/2757 Deng L.C.: J/A+A/462/535 Deng X.-F.: J/A+A/474/783 Deng Z.: J/AJ/117/2757 Deng Z.G.: J/A+AS/127/521 J/A+AS/132/181 J/AJ/112/628 J/ApJ/625/78 Denisenko D.V.: IX/36 Denisjuk O.: J/A+A/438/377 Denissenko D.V.: J/PAZh/30/692 Denisyuk E.K.: VI/93 Dennefeld M.: III/179 J/A+A/372/364 J/A+A/456/451 J/A+AS/116/95 J/A+AS/121/407 J/A+AS/126/297

J/A+AS/129/87 Dennerl K.: IX/10A IX/29 J/A+AS/127/119 Denninghoff S.: J/A+A/390/39 J/A+A/406/L9 J/A+A/410/813 Dennis J.W.III: J/BaltA/14/41 DENNISON B.: VIII/26 Denoyelle J.: II/20 III/154 Denzau H.: J/A+AS/125/399 Deo R.P.: J/AJ/132/321 Depagne E.: J/A+A/403/1105 J/A+A/416/1117 Depoy D.L.: J/AJ/109/709 J/AJ/130/1916 J/ApJ/470/864 J/ApJS/143/73 Deragopian G.: J/MNRAS/367/1609 Dere K.P.: J/A+A/466/771 J/A+AS/125/149 Derekas A.: J/A+A/394/943 J/A+A/394/97 J/A+A/410/879 J/A+A/461/613 J/A+A/463/589 DERMAN E.: J/A+AS/106/373 Derman E.: J/A+AS/125/407 Deroo P.: J/A+A/463/243 Derriere S.: J/ApJ/616/1284 Derry P.: J/A+A/476/1191 Derue F.: J/A+A/389/149 Dervisoglu A.: J/MNRAS/373/435 Desai V.: J/A+A/444/365 J/MNRAS/351/265 Desbats J.-M.: I/272 Desbats J.M.: J/A+AS/125/399 Descamps P.: J/A+A/425/1107 J/A+AS/125/399 Desert F.-X.: J/A+A/372/364 J/AJ/123/2070 Desert F.X.: J/ApJ/517/148 Desert J.-M.: J/A+A/385/986 Desiati P.: J/other/APh/26.282

Desidera S.: J/A+A/377/123 J/A+A/420/683 J/A+A/434/1107 J/A+A/454/581 J/A+A/464/953

J/A+A/471/515 J/AJ/121/327 Desmurs J.-F.: J/A+A/325/255 J/A+A/394/975 Despotes J. T.: J/ApJS/121/171 Desroches L.-B.: J/ApJ/632/799 Dessart L.: J/AJ/131/1648 Dessauges-Zavadsky M.: J/A+A/355/89 J/A+A/370/426 J/A+A/445/93 J/PASP/118/1077 Detal A.: J/A+AS/112/125 J/MNRAS/382/279 Detournay S.: J/A+A/399/1167 Dettbarn C.: I/264 I/286 J/A+AS/146/251 Dettmar R.J.: J/A+A/441/905 Dettmar R.-J.: J/A+A/458/341 J/A+AS/145/405 Deul E.: J/A+AS/137/75 J/A+AS/137/83 Deul E.R.: J/A+A/229/362 Deustua S.E.: J/A+A/437/789 Deutsch E.W.: J/AJ/124/1651 Deutsch L. K.: J/ApJS/122/221 Deutsch L.K.: J/ApJS/124/195 DEUTSCHMAN W.A.: II/6 Deveny J.B.: VII/11 Devereux N.: J/MNRAS/313/347 Devey C.R.: J/MNRAS/335/291 J/MNRAS/346/1143 J/MNRAS/351/1401 Deville Y.: J/A+A/469/575 Devine D.: J/ApJ/478/603 Devor J.: J/AJ/129/2856 J/ApJ/628/411 DEWDNEY P.E.: J/A+AS/105/385 Dewdney P.E.: J/ApJ/626/195 J/PASJ/52/997

Dey A.: J/AJ/112/2408 J/AJ/122/598 J/AJ/123/2223 J/AJ/127/213 J/ApJ/600/L99 J/ApJ/639/816

J/ApJ/641/140 J/ApJS/161/9 Deyoung T.: J/other/APh/26.282 Dezalay J.-P.: IX/36 J/other/GCN/1070.1 Dhillon V.S.: J/A+A/466/1031 Di Benedetto G. P.: J/A+A/339/858 Di Fabrizio: J/A+A/421/937 Di Fabrizio L.: J/A+A/430/603 Di Francesco J.: J/AJ/126/1916 di Francesco J.: J/ApJ/566/993 Di Franco S.: J/A+A/368/845 di Franco S.: J/A+AS/103/541 D.I.: J/A+A/394/395 Di Marcantonio P.: J/ApJ/601/485 J/ApJ/613/312 J/ApJ/634/1319 DI NELLA H.: J/A+AS/113/151 di Nella H.: J/MNRAS/283/367 Di Nella H.: J/MNRAS/287/472 Di Nella-Courtois H.: J/A+AS/135/133 Di Paola A.: J/AJ/129/2714 J/PAZh/30/243 DI PAOLANTONIO A.: J/A+AS/101/195 di Serego Alighieri S.: J/A+A/465/1099 di Stefano R.: J/ApJ/577/738 Di Stefano R.: J/ApJ/599/1067 J/ApJ/609/735 J/ApJ/610/247 Diaferio A.: J/AJ/120/2338 J/AJ/122/1289 J/AJ/124/1266 J/AJ/126/2152 J/AJ/128/1078 Diamond P.J.: J/A+A/391/967 J/A+A/476/L17 Diana S.: J/A+A/453/517 J/A+A/469/459 Dias J.E.: J/A+A/457/501 Dias W.S.: B/ocl J/A+A/376/441 J/A+A/388/168 J/A+A/404/913 J/A+A/446/949 DIAZ A.I.: J/A+AS/112/35 Diaz A.I.: J/AJ/110/212 J/MNRAS/358/521 J/MNRAS/361/1063 J/MNRAS/365/454 Diaz R.F.: J/A+A/469/309 Diaz R.J.: J/AJ/128/62 DIAZ S.: J/A+AS/104/557 DIAZ-CORDOVES J.: J/A+AS/110/329 Diaz-Cordoves J.: J/A+AS/114/247 DICKENS R.J.: J/MNRAS/250/314 Dickey J.: J/A+A/327/522 J/A+A/411/381 J/A+AS/116/193 Dickey J. M.: J/AJ/113/1939 Dickey J.M.: J/A+AS/126/325 J/AJ/96/1655 J/AJ/99/1071 J/ApJ/317/102 J/ApJ/473/822 J/ApJ/582/756

J/ApJ/626/887 J/ApJ/635/560 Dickinson C.: J/MNRAS/341/1066 J/MNRAS/360/340 Dickinson D.F.: J/ApJS/98/369 Dickinson M.: J/A+A/434/53 J/A+A/451/57 J/A+A/454/423 J/A+A/478/83 J/AJ/112/1335 J/AJ/120/2747

J/AJ/126/1183 J/AJ/127/131 J/AJ/127/180 J/AJ/127/2484 J/AJ/130/1 J/ApJ/480/568 J/ApJ/519/1 J/ApJ/551/111 J/ApJ/569/689 J/ApJ/592/728 J/ApJ/600/L99 J/ApJ/607/665 J/ApJ/613/200 J/ApJ/639/816 J/ApJ/640/603 J/ApJS/142/153 J/MNRAS/358/149

Dickinson M.E.: J/AJ/119/486 J/ApJ/606/L25 Dickman R.L.: J/ApJ/315/104 Dickson R.C.: J/A+A/472/805 Didelon P.: I/197A II/134 Dieball A.: J/A+A/376/745 J/A+A/391/547 J/A+AS/123/455 J/ApJ/625/156 Dieckvoss W.: I/61B I/69 J/AN/283/109 Diehl R.: J/A+AS/143/145 Diethelm R.: J/other/IBVS/5438 J/other/IBVS/5543 Dietrich: J/A+A/394/395 Dietrich J.P.: J/A+A/441/905 J/A+A/456/881 J/A+A/470/821 Dietrich M.: J/A+A/314/43 J/A+A/371/79 J/ApJ/510/659 J/ApJS/115/185 Dietrick J.: J/ApJ/612/511 Digby A.P.: J/other/Sci/292.698

Dimitrijevic M.S.: J/A+A/365/656 J/A+A/399/795 J/A+A/400/791 J/A+A/417/375 J/A+A/423/397

J/A+A/425/1147 J/A+A/434/773 J/A+A/441/391 DIMITRIJEVIC M.S.: J/A+AS/103/57 Dimitrijevic M.S.: J/A+AS/105/243 J/A+AS/105/245 J/A+AS/107/349 J/A+AS/109/551 J/A+AS/115/351

J/A+AS/116/359 J/A+AS/117/127 J/A+AS/119/369 J/A+AS/119/529 J/A+AS/120/373 J/A+AS/122/163 J/A+AS/122/533 J/A+AS/127/295 J/A+AS/127/543 J/A+AS/128/203 J/A+AS/128/359 J/A+AS/129/155 J/A+AS/130/539 J/A+AS/131/141 J/A+AS/131/143 J/A+AS/133/227 J/A+AS/140/191 J/A+AS/140/193

Dimitrov D.: J/A+A/467/229 J/other/PASA/23.154 DING M.D.: J/A+AS/110/99 Ding Y.-R.: J/A+AS/131/115 Dingus B.L.: J/ApJS/101/259 J/ApJS/107/227 J/ApJS/123/79 J/ApJS/94/551 Dintrans B.: J/A+A/475/1053 Dionatos O.: J/A+A/472/1055 Dionne M.: J/other/SerAJ/166.13Short Dipaolantonio A.: J/A+AS/125/399 Diplas A.: J/ApJS/93/211 Dirsch B.: J/A+A/433/43 J/A+A/459/391 J/AJ/127/2114 J/MNRAS/367/156 Disney M.J.: J/A+A/287/769 J/A+A/288/413 J/AJ/123/873 J/AJ/124/1954 J/AJ/124/690 J/AJ/128/16

J/MNRAS/337/641 J/MNRAS/342/738 J/MNRAS/346/787 J/MNRAS/355/1303 Distefano E.: J/A+A/469/503 Divan L.: III/187 Dixon K.L.: J/MNRAS/239/459 Dixon R.S.: VII/2A Dixon W.V.: J/AJ/104/495 J/ApJS/146/1 Djamaluddin T.: J/ApJS/89/57 Djenize S.: J/A+AS/102/607 Djorgovski S.: J/AJ/109/218 J/ApJ/409/28 J/PASP/97/932 Djorgovski S.G.: J/A+A/391/945 J/AJ/125/1660 J/AJ/125/2064 J/AJ/128/1017 J/AJ/130/425 J/ApJ/612/652 Djupvik A.A.: J/A+A/458/789

Djurasevic G.: J/A+A/343/894 J/A+A/374/638 J/A+A/376/158 J/A+A/415/283 J/A+A/420/1039

J/other/PASA/23.154 DJURIC Z.: J/A+AS/103/57 D.K.: J/A+A/385/1014 D.L.: J/A+A/465/271 D.N.: J/AZh/77/420 do Nascimento J.D.Jr: J/A+A/363/239 J/A+A/458/895 Dobashi K.: J/ApJS/110/21 J/ApJS/95/419 VII/244A Dobbie P.D.: J/MNRAS/342/1241 J/MNRAS/365/447 J/MNRAS/380/712 Dobbie P.J.K.: VII/79 Dobrotka A.: J/A+A/448/1107 Dobrzycka D.: J/ApJS/140/143 Dobrzycki A.: J/ApJ/457/102 J/ApJ/571/665 J/ApJS/130/37 J/ApJS/140/143 J/MNRAS/356/331 Dobson C.A.: J/AJ/107/751 Docobo J.A.: J/A+A/366/868 J/A+A/416/811 J/A+AS/130/117 J/AJ/127/1181 J/AJ/131/2681 Dodd R.J.: J/MNRAS/330/137 J/MNRAS/331/609 J/MNRAS/355/959 Dodelson S.: J/ApJ/596/437 Dodonov J.N.: J/other/AISAO/32.31 Dodonov S.N.: J/A+A/470/605 J/AZh/82/420

D’Odorico S.: J/A+A/318/347 J/A+A/341/641 J/A+A/370/426 J/A+A/373/757 J/A+A/375/1

J/A+A/392/395 J/A+A/437/883 D’odorico S.: J/A+A/445/93 D’Odorico S.: J/A+AS/127/217 J/AJ/120/2206 J/AJ/122/2190 D’odorico S.: J/MNRAS/335/555 D’Odorico V.: J/A+AS/127/217 J/AJ/119/2540 Dodson B.: J/AJ/129/656 Dodson R.: J/ApJ/616/110 J/PASJ/52/997 Dodson R.G.: J/AJ/127/3609 J/ApJS/155/33 Doebereiner S.: J/A+A/378/30 Doggett J.: J/AJ/110/2739 Doherty M.: J/MNRAS/361/525 Dohm-Palmer R.: J/AJ/120/2496 Dohm-Palmer R.C.: J/AJ/123/3154 J/AJ/125/1261 J/AJ/125/2502

Doi M.: J/AJ/121/2308 J/AJ/121/2331 J/AJ/121/31 J/AJ/122/503 J/AJ/123/567 J/AJ/125/1817 J/AJ/126/2579 J/ApJ/648/7 J/ApJS/137/279 J/ApJS/97/77 J/MNRAS/357/1357 J/PASJ/55/1079 J/PASJ/57/165 J/PASJ/58/313

Doi T.: J/AJ/124/445 J/AJ/125/277 J/AJ/127/3456 Doidinis S.P.: J/PASP/103/973 Doinidis S.P.: J/AJ/103/267 J/ApJS/103/433 J/PASP/102/1392 Dolan C.J.: J/AJ/118/2409 J/AJ/121/2124 J/AJ/123/387 Dolan J.F.: IX/6 Dolci M.: J/AJ/129/2714 Dolcini A.: J/A+A/469/503 Dole H.: J/A+A/372/364 J/AJ/131/250 J/ApJ/640/167 J/ApJS/154/155 Dolphin A.E.: J/A+A/396/473 J/A+A/398/479 J/AJ/116/1275 J/AJ/123/3154 J/AJ/125/1261 J/ApJ/521/577

J/ApJ/562/303 J/ApJS/166/549 J/PASP/117/37 Domingo A: J/A+A/343/446 Dominguez R.: J/A+AS/141/113 Dominici T.P.: J/A+A/409/857 J/A+AS/136/261 J/A+AS/139/321 Dominik C.: J/A+A/348/897 J/A+A/365/545 Dominik M.: J/A+A/313/841 Dommanget J.: I/196 I/211 I/260 I/274 Donahue M.: J/ApJ/569/689 J/ApJ/579/93 J/ApJ/616/1284 Donahue R.A.: J/AJ/111/439 Donas J.: J/A+A/303/661 J/A+A/327/522 J/AJ/130/825 J/ApJS/164/38 J/MNRAS/300/303 J/MNRAS/312/442 Donati A.: J/A+A/406/867 J/A+A/430/411 Donati J.F.: J/A+A/475/1053 Donato D.: J/A+A/375/739 J/A+A/433/1163 Dong F.: J/ApJS/148/243 Dong H.: J/MNRAS/371/38 Dong X.Y.: J/A+AS/139/575 J/AJ/131/1236 Donley J.L.: J/AJ/129/220 J/ApJ/642/126 Donnelly R.H.: J/AJ/131/527 J/ApJ/562/254 D’ONOFRIO M.: J/A+A/296/319 D’Onofrio M.: J/A+A/387/26 D’onofrio M.: J/A+A/470/39 D’ONOFRIO M.: J/A+AS/106/199 Donovan D.: J/ApJ/602/571 Donzelli C.: J/AJ/131/2035 Donzelli C.J.: J/AJ/129/1237 Dopita M.A.: J/AJ/126/2237 J/ApJ/449/592 J/ApJ/515/610 J/ApJ/619/755 J/ApJS/132/37 J/ApJS/165/307

J/ApJS/75/407 J/ApJS/76/1085 J/ApJS/83/87 Doppmann G.W.: J/AJ/130/1145 Dore L.: J/A+A/413/1177 J/A+A/455/1161 Dorfi E.A.: J/A+A/322/817 Dorman B.: J/A+A/324/915 J/AJ/108/1350 J/ApJ/522/983 VI/54 VI/65 Dorodhenko V.T.: J/PAZh/25/403 Doroshenko V.T.: III/208 J/A+A/376/775 J/A+A/422/925 J/A+A/437/87 J/ApJ/622/129 J/AZh/71/200

J/other/Ap/48.304 J/PAZh/21/652 J/PAZh/23/409 J/PAZh/26/535 J/PAZh/27/83 J/PAZh/27/942 Dorschner J.: III/95 J/A+A/449/583 dos Anjos S.: J/A+AS/125/329 Doss N.: J/MNRAS/357/471 Dotson J.L.: J/ApJ/604/717 Dottori H.: J/A+AS/128/545 J/AJ/116/723 Doty J.: J/other/GCN/1070.1 Doucet C.: J/A+A/402/745 Dougados C.: J/A+A/349/619 J/A+A/409/169 Dougados C: J/A+A/446/485 Dougados C.: J/A+A/463/1017 J/AJ/114/198 DOUGHERTY S.M.: J/A+A/290/609 Dougherty S.M.: J/PASJ/52/997 Douglas J.N.: VIII/42 Douglas K.: VIII/40 Douglas N.G.: J/MNRAS/369/120 J/MNRAS/369/97 Douglass G.C.: II/4 DOUGLASS G.G.: I/107A Douglass G.G.: I/157 I/237 J/AJ/117/1890 J/AJ/117/1905 J/AJ/117/2511 J/AJ/118/1395 J/AJ/119/3071 Dourneau G.: J/A+A/454/379 J/MNRAS/376/1707 Dowell C.D.: J/ApJ/604/717 DOWNES D.: J/A+A/293/703 Downes D.: J/A+AS/35/23 Downes J.J.: J/AJ/129/907 Downes R.A.: J/AJ/127/2838 V/110 V/123A V/79 V/94 Doxsey R. Golombek D.: VII/149 Doxsey R.: J/ApJ/409/28 Doyle M.T.: J/ApJS/165/307 J/MNRAS/368/651 Doyon R.: J/MNRAS/317/234 Draine B.T.: J/ApJ/633/871 Drake A.: J/ApJ/542/257 J/ApJ/598/597 Drake A.J.: J/AJ/127/334 J/AJ/131/1044 J/ApJ/552/289 J/ApJ/562/337 Drake C.L.: J/AJ/126/2237 Drake J.J.: J/AJ/108/1843 J/ApJS/150/19 J/ApJS/93/569 Drake N.A.: J/A+A/401/531 Drake S.A.: J/ApJ/616/1284 Drapatz S.: J/ApJ/466/254 Dravins D.: J/A+A/381/446 Drechsel H.: J/A+A/378/556 J/A+A/403/675 Dreizler S.: J/A+A/400/939 Dressel L.L.: J/ApJS/36/53 Dressler A.: J/A+A/470/39 J/A+AS/124/1 J/AJ/129/2138 J/AJ/95/284 DRESSLER A.: J/ApJ/368/54 Dressler A.: J/ApJ/591/53 J/ApJ/634/977 J/ApJS/110/213 J/ApJS/122/51 J/ApJS/75/241 J/ApJS/78/1

J/MNRAS/274/1107 VII/174 VII/176A Drew J.E.: J/MNRAS/341/805 J/MNRAS/362/689 Dreyer J.: J/other/APh/26.282 DREYER J.L.E.: VII/118 Driebe T.: J/A+A/339/123 J/A+A/466/1099 J/A+A/470/191 DRILLING J.S.: II/154 Drilling J.S.: III/137 J/AJ/73/590 J/PASP/107/846 DRILLING J.S.: VI/19 Drinkwater M.: J/MNRAS/372/425 J/MNRAS/375/68 Drinkwater M.J.: J/A+A/476/59 J/AJ/106/848 J/AJ/123/873 J/AJ/124/1954 J/AJ/124/690 J/AJ/128/16

J/AJ/130/957 J/AJ/131/1280 J/ApJS/165/307 J/MNRAS/279/595 J/MNRAS/337/641 J/MNRAS/344/188 J/other/PASA/16.113 VIII/34

Drira I.: J/A+AS/141/297 Drissen L.: J/AJ/118/1684 J/ApJS/93/187 Driver S.: J/MNRAS/372/425

Driver S.P.: J/AJ/130/1516 J/ApJ/453/48 J/MNRAS/329/227 J/MNRAS/329/87 J/MNRAS/333/100 J/MNRAS/350/1485 J/MNRAS/352/939 VII/240 VII/240 Droege T.F.: II/230 II/271A Drolias B.: J/MNRAS/351/1290 Dror D.H.A.: J/AJ/129/809 Drory N.: J/A+A/398/49 J/A+A/421/41 Drouin B.J.: J/ApJS/158/188 Drozdovsky I.: J/A+AS/124/559 J/A+AS/142/347 J/AJ/130/1324 Drozdovsky I.O.: J/A+AS/131/1 J/A+AS/142/425 Drukier G. A.: J/AJ/115/708 Drukier G.A.: J/A+A/375/L27 Drummond R.: J/A+A/463/243 Dryomova G.N.: J/A+A/437/375 Du B.T.: J/A+A/374/1056 Dubath P.: J/A+A/310/31 J/A+A/310/8 J/A+A/324/505 J/ApJ/635/214 Dubau J.: J/A+A/376/1113 J/A+AS/105/77 Dubernet L.: J/A+A/472/1029 Dubois P.: J/A+AS/91/371 VII/237 Dubovsky P.A.: J/other/IBVS/5586 Dubreuil D.: J/A+AS/124/283 Duc P.-A.: J/A+A/358/819 J/A+A/382/60 Duc P.A.: J/A+AS/129/455 Duc P.-A.: J/MNRAS/382/279 Ducati J.R.: II/237 II/66 Duchene G.: J/A+A/375/989 J/A+A/458/461 J/ApJ/636/932

Ducourant C.: I/272 I/300 I/303 J/A+A/361/1143 J/A+A/375/308 J/A+A/383/1054 J/A+A/409/361

J/A+A/437/1135 J/A+A/438/769 J/A+A/472/1055 Dudik R.P.: J/A+A/414/825 Dudinov V.N.: J/A+A/420/447

Duemmler R.: J/A+A/341/553 J/A+A/370/974 J/A+A/395/885 J/A+A/402/745 J/A+AS/123/209

J/A+AS/125/71 Duerbeck H. W.: III/174 DUERBECK H.W.: II/173 Duerbeck H.W.: J/A+A/418/249 J/A+AS/124/291 V/110 V/123A Duffard R.: J/A+A/468/L13 Duffee B.: J/A+A/373/625 Duffett-Smith P.J.: J/MNRAS/229/589 Duflot M.: III/190B III/192 DUFLOT M.: J/A+AS/110/177 Duflot M.: J/A+AS/124/255 J/A+AS/137/451 Dufour R.J.: J/AJ/104/92 J/AJ/130/1083 J/ApJ/513/168 Dufton P.L.: J/A+A/315/95 J/A+A/334/987 J/A+A/347/69 J/A+A/367/86 J/A+A/373/608 J/A+A/419/1123

J/A+A/434/1125 J/A+A/436/687 J/A+A/437/467 J/A+A/457/265 J/A+A/466/277 J/A+A/471/625 J/A+AS/139/231

Dufton Q.: J/AJ/122/2587 Duijsens M.F.J.: J/A+AS/119/37 Dullemond C.P.: J/A+A/409/1037 Dultzin-Hacyan D.: J/ApJ/639/37 J/ApJS/145/199 J/other/RMxAA/35.187 Dumitrescu A.: J/A+A/327/1017 J/A+AS/125/399 Dumont A.M.: J/A+A/376/775 Dumont A.-M.: J/A+A/422/925 Duncan D.K.: III/159A J/A+A/415/331 J/AJ/106/1059 J/AJ/119/3026 J/PASP/96/707 Duncan M.J.: J/AJ/116/1998 Duncan R.A.: J/other/PASAu/10.310 DUNHAM D.W.: I/125A Dunham D.W.: I/201 Dunlop J. S.: J/ApJ/511/612 Dunlop J.S.: J/MNRAS/238/1171 J/MNRAS/317/801 J/MNRAS/343/293 J/MNRAS/344/887 J/MNRAS/361/607

J/MNRAS/370/1057 J/MNRAS/372/1621 Dunn J.: J/ApJS/150/165 Dunne L.: J/MNRAS/372/1621 Duorah K.: J/A+A/469/713 Dupret M.-A.: J/A+A/415/241 Dupret M.A.: J/A+A/450/715 J/A+A/456/261 Dupret M.-A.: J/A+A/463/243 J/A+A/465/965 Duquennoy A.: J/A+A/282/831 J/A+A/314/846 DUQUENNOY A.: J/A+AS/108/25 Duquennoy A.: J/AJ/114/1087 DUQUENOY A.: III/96 Duran M.F.: J/ApJS/162/1 Durand N.: J/A+AS/102/57 Durand S.: J/A+AS/132/13 Durbala A.: J/A+A/449/937 Durdin J.M.: VIII/61 Durech J.: J/A+A/465/331 Durgapal A.K.: J/A+A/372/71 J/A+AS/122/111 Duric N.: J/A+A/435/437 J/AJ/128/1646 J/ApJS/120/247 Duriez L.: J/A+A/324/366 Durkee R.: J/A+A/465/331 Durlevich O.V.: B/gcvs II/172B II/205 II/219 II/250 J/other/IBVS/5721 J/PAZh/28/201 Durrance S.T.: J/AJ/121/2189 Durrell P.R.: J/ApJ/577/31 J/ApJ/582/170 J/ApJ/614/167 J/ApJS/145/65

Durret F.: J/A+A/393/439 J/A+A/416/475 J/A+A/429/101 J/A+A/451/1159 J/A+A/459/679 J/A+AS/100/583 J/A+AS/111/265 J/A+AS/122/409 J/A+AS/128/67 J/A+AS/129/281 J/A+AS/135/83 J/A+AS/139/525 J/A+AS/139/559

Duszlak J.: V/107 Dutra C.M.: J/A+A/359/347 J/A+A/383/631 J/A+A/397/177 J/A+A/404/223 J/AJ/117/238 J/AJ/119/1214

J/MNRAS/338/253 Dutra: J/A+A/400/533 Duus A.: VII/28 Duval D.: J/A+A/374/132 J/A+A/464/641 DUVENT J.L.: J/A+A/299/621 Duvoort M.R.: J/other/APh/26.282 Dvorak S.W.: J/other/IBVS/5502 J/other/IBVS/5542 Dwarakanath K.S.: J/AJ/118/625 J/other/JApA/25.143 Dwek E.: J/ApJ/459/686 Dwelly T.: J/MNRAS/362/1371 J/MNRAS/369/156 J/MNRAS/378/995 Dwivedi B.N.: J/A+A/375/591 DWORAK T.Z.: II/150A Dworetsky M.M.: J/A+A/271/515 J/A+A/274/335 Dye S.: II/253 J/A+A/377/442 J/MNRAS/379/1343 Dyl K.A.: J/ApJS/158/188 Dymock R.: J/A+A/465/331 Dzervitis U.: III/227 Dzhaniashvili I.B: J/AZh/78/408 Dziembowski W.A.: J/AcA/54/347 Dzura A.M.: J/A+AS/124/85 Dzyura A.M.: J/AZh/72/22

E.: J/A+A/397/177 J/A+A/400/533 J/A+A/409/541 J/A+A/421/937 J/A+A/421/937 J/A+A/445/939

J/A+AS/142/25 Eales S.: J/ApJ/517/148 Eales S.A.: J/MNRAS/350/785 J/MNRAS/372/1621 Earl J.: J/ApJ/448/179 Eastman J.D.: J/MNRAS/364/179 Eaton J.A.: J/AJ/109/1797 Eaton N.: J/MNRAS/351/1290 Ebbets D.C.: J/AJ/116/941

Ebeling H.: J/ApJ/580/774 J/ApJS/140/239 J/ApJS/140/265 J/MNRAS/274/75 J/MNRAS/281/799

J/MNRAS/301/881 J/MNRAS/306/857 J/MNRAS/318/333 J/MNRAS/330/1 J/MNRAS/339/913 Eberhardt K.: J/AJ/119/2801 Ebisawa K.: J/A+A/411/L59 J/ApJ/635/214 Ebrer J.: J/A+A/469/27 Ebrero J.: J/A+A/474/473 J/A+A/476/1191 Echevarria J.: J/A+A/275/201 Eck C.R.: J/ApJ/573/306 Eckart A.: J/A+A/433/117 J/ApJ/466/254 J/ApJ/523/248 Eckart M.E.: J/ApJ/596/944 J/ApJS/156/35 J/ApJS/165/19 Ecuvillon A.: J/A+A/449/723 J/A+A/458/997 Edelmann H.: J/A+A/400/939 J/A+A/442/1023 Edelson R.A.: J/A+A/369/57 Eder J.: J/ApJS/89/189 Eder J.A.: J/ApJS/131/47 J/MNRAS/359/1524 Ederoclite A.: J/A+A/459/875 Edge A.: J/MNRAS/372/425 J/MNRAS/375/68 Edge A.C.: J/A+A/425/367 J/MNRAS/274/75 J/MNRAS/281/799 J/MNRAS/301/881 J/MNRAS/306/857

J/MNRAS/318/333 J/MNRAS/327/588 J/MNRAS/330/1 J/MNRAS/339/913 J/MNRAS/366/645 J/MNRAS/372/1621 J/MNRAS/379/1343

Edge D.O.: VIII/1A Edmonds P.D.: J/ApJ/559/1060 J/ApJ/625/796 J/other/Sci/292.2290 Edvardsson B.: J/A+A/275/101 EDVARDSSON B.: J/A+A/293/75

Edvardsson B.: J/A+A/327/587 J/A+A/360/499 J/A+A/380/578 J/A+A/381/500 J/A+A/425/187

J/A+A/431/933 J/A+A/440/321 J/A+AS/102/603 EDVARSSON B.: II/160 Edwards A.L.: VIII/13 EDWARDS G.: J/MNRAS/247/182 Edwards P.G.: J/ApJ/616/110 J/ApJS/131/95 J/ApJS/155/33 J/PASJ/52/997 J/PASJ/55/351 Edwards R.T.: J/A+A/445/243 J/A+A/469/607 Edwards S.: III/240 Edwards W.R.: J/other/APh/26.282

Eenens P.: J/A+A/367/236 J/A+A/387/580 J/A+A/405/1087 J/A+A/408/611 J/A+A/455/1037

J/A+A/463/1061 J/A+A/464/263 Efimov Yu.S.: J/A+A/341/553 J/A+A/454/277 J/PAZh/26/535 Efremov Yu.N.: II/12 II/13 II/70 J/PAZh/21/13 Efremova B.V.: J/AJ/132/378 Efstahiou G.: J/MNRAS/289/263 Efstathiou A.: J/A+A/359/865 J/A+A/451/431 J/MNRAS/344/887 J/MNRAS/351/1290 Efstathiou G.: J/ApJS/107/201 J/MNRAS/246/433 J/MNRAS/269/151 J/MNRAS/308/897 J/MNRAS/329/227

J/MNRAS/329/87 J/MNRAS/333/100 J/MNRAS/350/1485 J/MNRAS/352/939 Efstathiou G.P.: J/MNRAS/279/459 VII/221 Egami E.: J/AJ/110/1993 J/ApJ/640/167 J/ApJ/642/126 J/ApJS/154/155 J/MNRAS/351/1290 Egan M.P.: III/231 J/A+A/453/535 J/AJ/122/1844 V/107 V/114 V/98 Eggen O.J.: J/AJ/107/2184 J/AJ/107/594 J/AJ/109/2229 J/AJ/110/823 J/AJ/115/2397 EGGEN O.J.: V/11 V/5 EGGEN O.J: V/8 Eggenberger P.: J/A+A/470/1009 Eggenkamp I.M.M.G.: II/202 J/A+A/304/415 Eggers D.: J/AJ/125/1696 J/AJ/127/1360 Eggleton P.P.: J/ApJ/629/1055 J/MNRAS/319/215 EGL ITIS I.: II/198 Eglitis I.: III/227 Egner S.: J/A+A/422/39 Egret D.: I/196 I/197A I/207 III/67A EGRET D.: III/68 Egret D.: III/83 IV/9 V/14 EGRET D.: V/16 Egret D.: V/17A EGRET D.: V/23 Ehgamberdiev S.: J/A+A/390/717 Ehle M.: J/A+A/426/11 J/A+A/448/1247 J/A+AS/141/469 Ehlerova S.: J/A+A/437/101 Ehrenfreund P.: J/A+A/429/559 Ehrlich R.: J/other/APh/26.282 EICHENDORF W.: VII/6 EICHHORN H.: I/73 Eichhorn H.: I/90 Eichler D.: J/ApJ/634/L89 Eigenbrod A.: J/A+A/423/189 Eikenberry S.S.: J/ApJ/635/280 Einasto J.: J/A+A/336/35 J/A+A/459/L1 J/A+A/462/811 J/A+AS/123/119 J/AJ/122/2222 J/AJ/123/51

J/AN/327/365 J/MNRAS/269/301 Einasto M.: J/A+A/336/35 J/A+A/459/L1 J/A+A/462/811 J/A+AS/123/119 J/AJ/122/2222 J/AJ/123/51

J/AN/327/365 J/MNRAS/269/301 Einicke O.H.: I/99 I/99 J/A+A/312/1031 Eiroa C.: J/A+A/310/143 J/A+A/365/110 J/A+A/378/116 J/A+A/379/564 J/A+A/380/609 J/A+A/401/613 Eisenhardt P.: J/AJ/123/2223 J/AJ/127/2484 Eisenhardt P.R.: J/ApJ/639/816 J/ApJS/142/153 Eisenhauer F.: J/ApJ/643/1011 Eisenstein D.: J/AJ/126/1023 J/AJ/126/2579 J/AJ/130/1 J/AJ/131/49 J/AJ/132/231 J/MNRAS/372/425

J/MNRAS/375/68 Eisenstein D.J.: J/AJ/125/1817 J/AJ/131/1674 J/AJ/132/676 J/ApJ/606/L25 J/ApJ/607/426 Eisloeffel J.: J/A+A/416/555 J/A+A/419/249 Eisloffel J.: J/A+A/429/1007 Eissner W.: J/A+A/408/789 J/A+AS/137/165 EKBERG J. O.: J/A+AS/101/1 Eker Z.: J/MNRAS/349/1069 J/MNRAS/357/497 J/MNRAS/373/1483 EKER Z.: V/71 Ekers J.A.: J/MNRAS/371/898 Ekers R.D.: J/A+A/457/517 J/A+AS/106/1 J/AJ/104/704 J/AJ/123/873 J/AJ/124/1954 J/AJ/124/690

J/AJ/128/16 J/AJ/130/1358 J/MNRAS/337/641 J/MNRAS/352/1245 J/MNRAS/354/305 J/MNRAS/355/1303 J/MNRAS/371/898 J/other/SerAJ/166.13Short

EKERS R.D.: VIII/27 VIII/27 Ekers R.D.: VIII/38 VIII/38 VIII/38 VIII/38 VIII/63 Ekholm T.: J/A+A/383/398 J/A+A/393/57 Ekmekci F.: J/A+A/391/641 J/AN/323/31 El Eid M.F.: J/ApJ/611/452 Elbaz B.: J/A+A/372/364 Elbaz D.: J/A+A/342/313 J/A+A/382/60 J/A+A/407/791 J/A+A/423/867 J/A+A/427/23 J/A+A/427/773

J/A+A/430/115 J/A+A/451/57 J/A+A/462/81 J/A+A/465/1099 J/MNRAS/351/1290 Elfhag T.: J/A+AS/115/439 Elias II N.M.: J/AJ/114/2822 J/AJ/120/1548 Elias J.H.: J/AJ/112/2408 Elias N.M. II: J/ApJ/439/983 ELIAS N.M.: J/A+AS/102/343 Elias N.M.II: J/ApJ/484/394 Eliasdottir A.: J/A+A/455/L1 Elias-Rosa N.: J/MNRAS/369/1880 Eliche-Moral C.: J/ApJS/154/155 Ellinger Y.: J/A+A/319/318 Ellingsen S.: J/MNRAS/367/1609 Ellingsen S.P.: J/A+A/403/1095 J/AJ/127/3609 J/AJ/130/2529 J/MNRAS/359/1498 J/MNRAS/377/571 Ellingson E.: J/ApJ/471/694 J/ApJ/566/744 J/ApJ/636/631 J/ApJS/113/1 J/ApJS/116/211 J/ApJS/116/231

J/ApJS/116/247 J/ApJS/129/475 J/ApJS/157/228 Ellington E.: J/ApJS/102/289 Elliot J.L.: J/AJ/104/862 ELLIOT J.L.: J/AJ/107/1814 Elliot J.L.: J/AJ/109/1352 J/AJ/119/2008 J/AJ/119/936 Elliot K.H.: VII/232 Elliott K.H.: J/MNRAS/274/1194 Ellis R.: J/A+A/437/789 J/MNRAS/273/157 J/MNRAS/297/885 VII/23 ELLIS R.S.: II/178

Ellis R.S.: J/AJ/112/359 J/ApJ/497/188 J/ApJ/499/112 J/ApJ/551/111 J/ApJ/591/53 J/ApJ/627/32

J/ApJ/633/174 J/ApJ/634/977 J/ApJS/107/1 J/ApJS/110/213 J/ApJS/122/51 J/MNRAS/235/827 ELLIS R.S.: J/MNRAS/244/408 J/MNRAS/253/686 Ellis R.S.: J/MNRAS/300/303 J/MNRAS/308/897 J/MNRAS/311/456 J/MNRAS/312/442 J/MNRAS/329/227

J/MNRAS/329/87 J/MNRAS/330/75 J/MNRAS/333/100 J/MNRAS/350/1485 J/MNRAS/352/939 J/MNRAS/361/525 VII/246

Ellis S.C.: J/MNRAS/367/627 Ellison S.L.: J/AJ/130/1345 J/PASP/118/1077 VIII/34 Ellman N.: J/AJ/127/1158 J/AJ/127/1811 VII/200 Ellsworth R.W.: J/other/APh/26.282 ELMEGREEN B.G.: J/AJ/107/1453 Elmegreen B.G.: J/AJ/121/182 J/AJ/128/2170 J/ApJ/631/85 J/ApJS/162/49 ELMEGREEN D.M.: J/AJ/107/1453 Elmegreen D.M.: J/AJ/121/182 J/ApJ/631/85 Elsaesser H.: J/A+AS/109/537 J/A+AS/116/43 J/A+AS/117/261 ELSMORE B.: I/169 Elsner R.F.: J/ApJ/615/897 Elston R.: J/AJ/112/2408 J/AJ/123/2223 Elston R.J.: J/ApJ/639/816 Elvis M.: J/ApJS/105/369 J/ApJS/95/1 J/MNRAS/344/134 VII/19 Elwell B.: J/PASP/111/438 E.M.: J/A+A/395/753 J/A+AS/142/25 Emelianov N.V.: J/A+AS/139/47 Emeliyanov V.A.: J/A+AS/132/233 Emelyanov A.N.: J/PAZh/26/355 Emelyanov N.: J/A+A/391/767 Emerson J.P.: J/A+A/365/90 Emery R.J.: II/56 Endesfelder U.: J/A+A/469/207 Endl M.: J/A+A/420/683 Engargiola G.: J/ApJS/149/343 Engel E.A.: J/MNRAS/357/471 Engelbracht C.W.: J/AJ/131/250 J/ApJ/633/871 J/ApJS/151/387 J/ApJS/154/155 Engelke C.W.: J/AJ/132/1445 J/ApJS/150/343 V/114 ENGELS D.: J/A+A/296/727 Engels D.: J/A+A/322/159 J/A+A/366/771 J/A+A/374/907 J/A+A/397/463 J/A+A/400/939 J/A+A/401/927 J/A+A/406/535 J/A+A/414/487 J/A+A/431/779 J/A+A/442/109 J/A+A/446/773 J/A+A/449/425 J/A+A/458/533

ENGELS D.: J/A+AS/110/469

Engels D.: J/A+AS/127/145 J/A+AS/128/507 J/A+AS/134/483 J/A+AS/135/511 J/A+AS/137/299

J/A+AS/142/417 ENGELS D.: V/18 Engin S.: J/A+A/371/1078 Engle S.: J/A+A/463/1061 Englhauser J.: IX/10A IX/29 J/A+AS/127/145 English J.: J/AJ/118/752 Enoch B.: J/A+A/467/785 J/MNRAS/372/1117 J/MNRAS/379/647 J/MNRAS/379/816 Enoch M.L.: J/ApJ/638/293 Enya K.: J/ApJS/141/23 J/ApJS/141/31 EPCHTEIN N.: II/177 Epchtein N.: J/A+A/227/82 EPCHTEIN N.: J/A+A/299/453 Epchtein N.: J/A+A/334/609 J/A+A/349/L69 J/A+A/363/62 J/A+A/380/590 J/A+A/387/1 J/A+A/401/959

J/A+A/413/1037 J/A+AS/122/489 J/A+AS/135/133 EPCHTEIN N.: J/A+AS/99/31 Epchtein N.: J/ApJ/552/289 VII/242 Epps E.A.: V/32A V/4 Epremian R.A.: III/253 Eracleous M.: J/ApJ/599/886 J/ApJ/613/129 J/ApJS/150/181 Erastova L.K.: J/A+AS/147/169 ERASTOVA L.K.: VII/172 Erben T.: J/A+A/441/905 J/A+A/462/875 J/A+A/470/821 J/A+AS/137/75 J/A+AS/137/83 Erdem A.: J/A+A/387/240 J/A+A/417/263 Erdogdu P.: J/MNRAS/352/939 Erickson E.F.: J/ApJS/162/346 Erickson W.C.: VIII/79 Eriksen K.A.: J/AJ/131/527 Erikson A.: J/A+AS/113/115 Eriksson K.: J/A+A/350/955 ERIKSSON P-I.W.: II/120 Eritsian M.A.: J/other/Ap/47.305 Erkapic S.: J/A+A/343/894 J/A+A/374/638 J/A+A/376/158 J/A+A/415/283 J/A+A/420/1039 Erkens U.: J/A+A/323/707 Ermakov S.V.: J/A+A/397/1035 Errico L.: J/A+A/367/199 J/A+AS/114/341 J/A+AS/136/95 Erspamer D.: J/A+A/398/1121 Ertan A.Y.: J/A+AS/129/9 Erwin P.: J/ApJS/154/155 ESA 1997: I/239 Escalera E.: J/A+A/310/31 J/A+A/310/8 J/A+AS/117/519 Escudero A.V.: J/A+A/380/300 J/A+A/427/231 Eshankulova M.: J/A+A/374/638 Esipov V.F.: J/PAZh/25/30 J/PAZh/27/841 J/PAZh/29/858 J/PAZh/30/297 J/PAZh/30/855 J/PAZh/32/48

J/PAZh/32/662 J/PAZh/32/737 J/PAZh/33/679 Esko E.: J/A+A/383/197 Eskridge P.B.: J/ApJ/442/523 J/ApJS/143/73 J/ApJS/97/141 Esmond J.R.: VI/61 ESO: II/234 Espada D.: J/A+A/436/443 J/A+A/449/937 J/A+A/462/507 J/A+A/470/505 J/A+A/472/121 Espenak F.: VI/97 Espey B.R.: J/AJ/119/486 J/ApJ/636/1002 Esposito J.A.: J/ApJS/101/259 J/ApJS/107/227 J/ApJS/123/79 J/ApJS/94/551 Estalella R.: J/A+A/474/911 J/ApJ/463/205 Esteban C.: J/ApJ/595/247 J/ApJS/120/113 J/MNRAS/362/301 J/MNRAS/368/253 Etherton J.: J/A+A/414/23 Etienne A.: J/ApJS/101/259 J/ApJS/107/227 Ettori S.: J/A+A/434/53 J/MNRAS/276/689 ETZEL P.B.: J/A+A/291/110 Etzel P.B.: J/AJ/109/1269 J/AJ/109/1308 J/MNRAS/363/529 Eubanks M.T.: J/AJ/110/880 Eubanks T.M.: J/A+A/375/661 European Southern Observatory (ESO): B/eso

Evans A.: J/A+A/384/504 J/A+A/412/405 J/A+AS/131/401 J/MNRAS/372/1117 J/MNRAS/379/647

J/MNRAS/379/816 Evans A.S.: J/AJ/121/1840

Evans C.J.: J/A+A/437/467 J/A+A/456/623 J/A+A/457/265 J/A+A/466/277 J/A+A/471/625

J/MNRAS/353/601 Evans D.S.: I/110 III/47B Evans D.W.: I/282 J/A+A/330/515 J/A+AS/130/157 Evans Ii N.J.: J/ApJ/638/293 Evans I.N.: J/ApJS/105/93 J/ApJS/143/257 Evans N.R.: J/ApJ/589/509 J/ApJS/150/19 Evans N.W.: J/AJ/131/895 J/MNRAS/330/792 J/MNRAS/351/1071 J/MNRAS/352/285 J/MNRAS/369/120

J/MNRAS/369/97 Evans P.A.: J/MNRAS/367/781 Evans R.: J/MNRAS/352/478 Evans R.W.: J/AJ/117/1700 Evans T.: II/246 Evenson P.A.: J/other/APh/26.282 Everett M.E.: J/PASP/114/656 Evrard A.E.: J/AJ/130/968 EVREN S.: J/A+A/281/811 J/A+A/287/817 Evren S.: J/A+A/336/920 J/A+A/374/1056 J/A+A/388/298 J/A+AS/129/9 J/MNRAS/376/573 Evstigneeva N.M.: J/PAZh/20/693 J/PAZh/21/509 J/PAZh/23/454 EXOSAT Observatory: VI/36 VI/41 Exter K.: J/A+A/452/503

Eyer L.: J/AcA/52/241 J/MNRAS/327/601 J/MNRAS/331/45 J/MNRAS/340/1346 J/MNRAS/349/1059

J/MNRAS/356/331 J/MNRAS/358/30 Ezhkova O.: J/A+AS/124/33

Ezhkova O.V.: J/A+A/350/513 J/A+AS/139/513 J/AZh/78/408 J/MNRAS/310/210 J/PAZh/27/83 J/PAZh/27/942

F.: J/A+A/384/371 J/A+AS/136/509 J/A+AS/139/483 J/A+AS/31/439 J/MNRAS/326/959 F. Perez Hernandez F. Fernandez A.: J/A+A/468/637 Fabbian D.: J/A+A/434/235 Fabbiano G.: J/ApJ/442/523 J/ApJ/586/826 J/ApJS/161/271 J/ApJS/80/531 J/ApJS/97/141 J/MNRAS/344/134 Fabbro S.: J/A+A/437/789 FABER S.M.: III/22 Faber S.M.: J/AJ/114/1365 J/AJ/127/3121 J/AJ/129/2138 FABER S.M.: J/ApJ/368/54 Faber S.M.: J/ApJ/481/673 J/ApJ/571/136 J/ApJ/599/1006 J/ApJ/625/6 J/ApJS/109/333 J/ApJS/116/1

J/ApJS/142/1 J/ApJS/157/175 J/ApJS/159/41 FABER S.M.: J/ApJS/94/687 Faber S.M.: J/MNRAS/371/221 VII/176A VII/198 Fabian A.C.: J/A+A/444/79 J/A+A/476/1191 J/MNRAS/274/75 J/MNRAS/301/881 J/MNRAS/306/857

J/MNRAS/318/333 Fabian D.: J/AJ/111/645 Fabiani Bendicho P.: J/A+A/423/1109

Fabregat J.: J/A+A/329/579 J/A+A/394/479 J/A+A/472/565 J/A+A/472/577 J/A+AS/119/271

J/A+AS/123/1 J/A+AS/123/59 J/A+AS/128/139 J/MNRAS/358/66 Fabricant D.: J/AJ/105/788 J/ApJ/498/195 J/ApJ/500/714 J/ApJ/539/577 J/ApJ/641/140 J/ApJS/126/271

J/ApJS/126/331 J/ApJS/161/9 Fabricant D.G.: J/AJ/122/2893 J/AJ/122/714 J/AJ/129/363 J/AJ/131/527 J/ApJ/381/33 J/ApJ/470/724

J/ApJ/559/606 Fabricius C.: I/250 Fabricius C: I/259

Fabricius C.: I/275 I/276 I/99 I/99 IV/25 J/A+A/312/1031 J/A+A/356/141 J/A+A/372/95

J/A+A/373/576 J/A+AS/144/45 Fabrika S.: J/A+AS/140/309 Fabrika S.N.: J/PAZh/25/18 Fadda D.: J/A+A/382/60 J/A+A/427/23 J/A+A/427/773 J/AJ/128/1 J/AJ/131/250 J/MNRAS/351/1290

J/MNRAS/358/397 J/MNRAS/371/1891 Fadeyev V.: J/A+A/437/789 Fagas M.: J/A+A/465/331 Fagiolini M.: J/A+A/462/107 Fagotto F.: J/A+A/311/361 J/A+A/333/479 J/A+AS/106/275 VI/96 Fahlamn G.G.: J/AJ/117/247 Fahlman G.C.: J/AJ/109/663 FAHLMAN G.G.: J/AJ/100/1811 J/AJ/103/460 Fahlman G.G.: J/AJ/109/650 J/AJ/114/1488 J/AJ/127/2771 J/ApJS/120/265 Faifer F.R.: J/MNRAS/355/608 Fairall A.P.: J/A+A/297/617 J/A+A/352/39 J/A+AS/127/463 J/AJ/116/1 J/MNRAS/366/267 VII/107

VII/79 Fairall F.P.: VII/142A Falanga M.: J/A+A/458/21 Falchi A.: J/A+A/326/249 FALCKE H.: J/A+A/298/375 Falcke H.: J/A+A/409/115 J/A+A/435/521 J/A+AS/135/273 Falco E.E.: J/AJ/124/601 J/AJ/130/1977 J/AJ/131/527 J/ApJ/560/566 J/PASP/111/438 Falcon-Barroso J.: J/MNRAS/371/703 J/MNRAS/374/664

Falgarone E.: J/A+A/350/883 J/A+A/357/219 J/A+A/372/173 J/A+A/420/945 J/A+A/421/623

J/A+A/469/173 VIII/66 VIII/66 Falin J.L.: J/A+AS/135/231 Fall S.M.: J/AJ/114/2381 J/ApJ/561/727 Fallon A.A.: J/ApJS/147/305 Fallon F.W.: I/92 Falomo R.: J/AJ/104/28 Famaey B.: J/A+A/430/165 FaN J.H.: J/A+A/278/6 Fan J.H.: J/A+A/369/758 J/A+A/407/899 J/A+A/436/799 FAN J.H.: J/A+AS/106/361

Fan X.: J/AJ/112/628 J/AJ/117/2757 J/AJ/121/2308 J/AJ/121/31 J/AJ/122/503 J/AJ/123/3409 J/AJ/123/567 J/AJ/124/1651 J/AJ/126/1023 J/AJ/126/2209 J/AJ/126/2579 J/AJ/127/3553 J/AJ/130/1 J/AJ/131/1 J/AJ/131/2722 J/AJ/131/49 J/AJ/132/231 J/ApJ/606/L25 J/ApJS/123/377 J/ApJS/155/257 J/ApJS/165/1

FAN X.-M.: J/AJ/106/426 Fan Y.: J/A+A/456/261 J/A+AS/127/181 J/other/PPMtO/14.185 Fan Z.: J/A+A/449/143 Fanelli M.N.: J/AJ/127/3213 FANELLI M.N.: J/ApJ/413/604 Fanelli M.N.: J/ApJS/104/287 J/ApJS/82/197 FANG C.: J/A+AS/110/99

Fang F.: II/255 J/A+A/467/73 J/AJ/129/1198 J/AJ/129/2074 J/AJ/131/250 J/MNRAS/371/1891 J/PASP/113/10 Fang Li-Z.: J/AJ/117/2757 Fang L.-Z.: J/A+A/338/813 Fang L.Z.: J/AJ/112/628 Fang L.-Z.: J/ApJS/115/185 Fang M.: J/other/ChA+A/28.308 Fang M.J.: J/A+A/395/587 J/A+AS/122/515 J/A+AS/136/515 FANSELOW J.: I/169 Fanti C.: J/A+A/369/380 J/A+A/379/755 J/A+A/426/463 J/A+A/446/97 J/A+A/462/449 J/A+AS/131/303

J/A+AS/43/1 Fanti et al: VIII/36 Fanti R.: J/A+A/363/887 J/A+A/369/380 J/A+A/423/935 J/A+A/426/463 J/A+A/439/487 J/A+A/446/97

J/A+A/462/449 J/A+AS/120/89 J/A+AS/131/303 Faraggiana R.: J/A+A/379/162 J/A+A/381/105 J/A+A/412/447 J/A+A/472/241 J/A+AS/137/273 Farcas L.: J/A+AS/125/399 Farfanelli F.: J/A+A/469/503 Faria D.: J/A+A/465/357 Farihi J.: J/ApJS/161/394 Farinas R.: J/PASP/108/332 Farinelli R.: J/A+A/458/21 Farisato G.: J/AJ/121/327 Farnsworth A.H.: II/221A Farquhar P.R.A.: J/A+AS/121/139 Farraggiana R.: J/A+AS/135/503 Farrah D.: J/A+A/467/565 J/AJ/129/1198 J/AJ/129/2074 J/MNRAS/372/1621 Farrell W.M.: J/ApJ/612/511 Fasano G.: J/A+A/364/517 J/A+A/387/26 J/A+A/470/39 J/A+AS/129/583 J/ApJ/542/673 Fassett C.I.: J/ApJ/529/723 Fassnacht C.D.: J/AJ/130/923 J/ApJ/600/L155 J/MNRAS/377/430 VIII/72 Fathi K.: J/A+A/414/23 J/A+A/466/905 Faucher P.: J/A+AS/105/77 Fauerbach M.: J/A+A/465/331 Faulkner A.: J/MNRAS/352/1439 Faulkner A.J.: J/MNRAS/342/1299 J/MNRAS/372/777 Faundez S.: J/A+A/426/97 FAUNDEZ-ABANS M.: J/A+AS/104/1 Faundez-Abans M.: J/A+AS/129/357 J/A+AS/130/455 Faure A.: J/A+A/472/1029 Faure C.: J/A+A/391/481 Fauvaud S.: J/A+A/465/331 Favata F.: J/A+A/345/521 J/A+A/355/651 J/A+A/399/919 J/A+A/404/927 J/A+A/412/495 J/A+A/416/281

J/A+A/417/945 J/A+A/426/1007 J/A+A/445/673 J/A+A/459/199 J/A+A/461/977 J/A+A/463/275 J/ApJS/160/450 J/MNRAS/326/1161

Fayet G.: I/36 Fazely A.R.: J/other/APh/26.282 Fazio G. Hrivnak B. J.: J/ApJS/122/221 Fazio G.: J/AJ/131/1574 J/ApJS/161/9 Fazio G.G.: J/A+A/403/493 J/ApJ/641/140 J/ApJ/642/126 J/ApJ/643/965 J/ApJS/154/155 J/ApJS/96/117 FEAST M.: J/MNRAS/267/711 Feast M.W.: J/MNRAS/121/337 J/MNRAS/317/460 J/MNRAS/369/751 J/MNRAS/369/783 J/MNRAS/370/1979

J/MNRAS/376/313 FECKEL F.C.: V/71 FEDELE S.B.: J/AJ/105/2299 Fedele S.B.: J/AJ/106/1059 FEDELE S.B.: J/AJ/106/1080 Federici L.: J/A+A/416/917 J/A+A/456/985 J/A+A/471/127 J/A+A/473/429 Federman S.R.: J/ApJS/151/313 Federspiel M.: J/A+A/327/952 J/A+AS/125/399 Fedorov P.: I/302 Fedorova A.A.: J/PAZh/23/90 Fedorovich V.P Frolov M.S.: II/140 FEHMERS G.C.: J/A+AS/108/61 FEHRENBACH C.: J/A+AS/110/177 Fehrenbach C.: J/A+AS/124/255 Fehrenbach Ch.: III/192 Feibelman W. A.: J/PASP/105/1279 Feibelman W.A.: J/A+AS/139/135 J/ApJ/614/745 Feigelson E.: J/A+AS/134/221

Feigelson E.D.: J/A+A/323/739 J/A+A/359/113 J/A+A/468/353 J/A+AS/122/235 J/AJ/120/1426 J/AJ/122/2810 J/AJ/131/2164 J/ApJ/508/314 J/ApJ/525/127 J/ApJ/554/742 J/ApJ/574/258 J/ApJ/575/354 J/ApJ/584/911 J/ApJ/589/225 J/ApJ/613/1179 J/ApJ/614/267 J/ApJS/160/319 J/ApJS/160/353 J/ApJS/160/450 J/ApJS/160/511 J/ApJS/163/306

FEIGELSON E.D.: J/ApJS/93/145 FEIJTH H.: II/151 Feiler R.: J/A+A/474/1093 FEINSTEIN A.: J/A+A/301/396 Feinstein A.: J/A+A/402/549 FEINSTEIN A.: J/A+AS/106/339 J/A+AS/111/85 Feinstein A.: J/A+AS/116/75 J/A+AS/123/63 J/A+AS/124/13 J/A+AS/137/233 Feinstein C.: J/A+A/409/933 J/A+A/430/471 J/A+A/462/621 J/A+A/471/165 J/ApJS/112/29 Feissel M.: I/251 J/A+A/375/661 J/A+AS/116/473 Feissel-Vernier M.: J/AJ/127/3587 FEITZINGER J.V.: III/93 VII/68A Fekel C.: J/A+AS/137/369 Fekel F.C.: B/sb9 J/A+A/334/863 FEKEL F.C.: J/AJ/105/2265 Fekel F.C.: J/AJ/129/1001 J/AJ/129/1669 J/AJ/129/2026 J/AJ/129/2815 J/AJ/130/794 V/122 V/76 Feldman F.R.: VII/171 Feldman P.A.: J/PASP/107/617 Feldman P.D.: J/AJ/128/1990 J/ApJ/612/202 J/ApJ/621/53 J/ApJS/150/1 Feldman U.: J/A+A/304/563 J/A+A/375/591 J/A+AS/126/281 Feldmeier J.J.: J/ApJ/479/231 J/ApJ/492/62 J/ApJ/577/31 J/ApJ/582/170 J/ApJ/614/167 J/ApJS/145/65 Feldt M.: J/ApJS/155/123 Felenbok P.: J/A+A/415/9 J/A+AS/129/281 J/A+AS/130/323 J/A+AS/139/525 J/ApJ/520/491 FELLI M.: J/A+A/284/233 J/A+A/291/261 Felli M.: J/A+A/362/199 J/A+A/368/845 J/A+A/392/971 J/A+A/453/911 J/A+AS/101/153 J/A+AS/103/541

J/ApJ/552/289 Feltzing S.: J/A+A/377/911 J/A+A/385/67 J/A+A/410/527 J/A+A/415/155 J/A+A/433/185 J/A+A/465/357

J/A+A/467/665 J/A+AS/129/237 Fender R.P.: J/MNRAS/298/692 J/MNRAS/308/473 Fenimore E.: J/other/GCN/1070.1 FERANCE S.: J/ApJ/412/183 FERETTI L: J/A+A/281/375 Feretti L.: J/A+A/370/409 J/A+A/416/839 J/A+AS/74/315 Ferguson C.: J/ApJS/156/69 Ferguson H.: J/AJ/112/2408 J/ApJ/613/200 Ferguson H.C.: J/A+A/434/53 J/A+A/454/423 J/A+A/478/83 J/AJ/104/495 J/AJ/112/1335 J/AJ/119/486

J/AJ/120/2747 J/AJ/127/131 J/AJ/127/2738 J/AJ/130/1693 J/ApJ/504/113 J/ApJ/532/308 J/ApJ/600/L155 J/ApJ/600/L99 J/ApJ/607/665 J/ApJS/165/307 J/MNRAS/277/125 VII/160 VII/180 VII/180

Ferland G.J.: J/A+A/416/383 J/A+A/424/363 J/A+A/433/745 J/A+A/455/1157 Ferlet R.: J/A+A/303/137 J/A+A/306/924 J/A+A/365/110 J/A+A/378/116 J/A+A/379/564 J/A+A/385/986

J/A+A/389/149 J/A+A/454/185 FERLET R.: J/A+AS/109/447 Ferlet R.: J/A+AS/145/275 FERLET R. NITSCHLEM C.: J/A+A/299/557 Fernandes R.C.: J/MNRAS/366/480 FERNANDEZ A.: VII/88 Fernandez E.: J/AJ/126/632 J/AJ/127/180 Fernandez J.A.: J/A+AS/146/73 Fernandez M.: J/A+A/310/143 J/A+A/349/619 J/A+A/353/1044 J/A+A/375/130 J/A+A/409/169 FERNANDEZ M.: J/A+AS/101/485 Fernandez M.: J/A+AS/113/473 Fernandez-Figueroa M.J.: J/A+A/312/221 J/A+A/326/780 J/A+A/379/976 J/A+AS/123/473 J/A+AS/146/103

J/MNRAS/328/45 Fernandez-Soto A.: J/A+A/469/503 J/ApJ/513/34 J/ApJ/538/493 J/ApJS/135/41 Fernie J.D.: J/A+A/275/484 J/A+A/394/617 FERNIE J.D.: J/AJ/105/2291 Fernley J.: J/A+A/330/515 J/A+AS/122/131 J/A+AS/125/313 J/A+AS/125/321 Feroci M.: J/A+A/455/813 J/A+AS/124/123 J/ApJS/156/217 Ferrara A.: J/A+A/434/801 Ferrarese L.: J/AJ/108/1598 J/ApJ/463/26 J/ApJ/507/655 J/ApJ/508/491 J/ApJ/512/48 J/ApJ/514/614

J/ApJ/521/155 J/ApJ/523/540 J/ApJ/529/723 J/ApJ/613/279 J/ApJ/613/682 J/ApJS/153/223 J/ApJS/154/509 Ferrari C.: J/A+A/399/813 J/A+A/430/19 J/A+A/461/81 FERRARO F.R.: J/A+A/290/69

Ferraro F.R.: J/A+A/298/461 J/A+A/320/757 J/A+A/324/915 J/A+A/419/139 J/A+A/420/173

J/A+A/462/599 J/A+A/464/201 FERRARO F.R.: J/A+AS/84/59 Ferraro F.R.: J/AJ/106/154 J/AJ/118/2839 J/AJ/121/3114 J/AJ/122/2569 J/AJ/124/3222 J/AJ/124/915

J/ApJ/522/983 J/ApJ/640/L43 J/ApJ/654/915 FERRARO F.R.: J/ApJS/85/293 J/MNRAS/252/357 Ferraro F.R.: J/MNRAS/264/273 FERRARO F.R.: J/MNRAS/272/391 Ferraro F.R.: J/MNRAS/349/1278 J/MNRAS/351/1204 J/MNRAS/354/708 J/MNRAS/357/265 J/MNRAS/360/185 FERRARO I.: J/A+AS/67/327 Ferraro: J/MNRAS/361/272 Ferreiro D.L.: J/A+A/428/837 Ferrer O.E.: J/A+A/343/841 Ferreras I.: J/ApJ/635/243 Ferrero L.: J/A+A/447/921 Ferrin I.: J/AJ/127/1158 Ferro A. Arellano: J/A+A/275/484 Feruglio C.: J/A+A/457/501 J/A+A/466/31 Fesen R.A.: J/ApJ/613/343 Feser T.: J/other/APh/26.282 Festin L.: J/A+AS/113/115 Fetisova T.S.: III/208 Feuchtgruber H.: J/A+A/390/1033 J/ApJ/566/880 Feuchtinger M.U.: J/A+A/322/817 Feulner G.: J/A+A/421/41

Fey A.L.: J/AJ/110/880 J/AJ/127/3587 J/AJ/127/3609 J/AJ/128/1570 J/AJ/130/2529 J/ApJ/614/607

J/ApJS/128/17 F.F.S.: J/A+A/432/369 Ficarra A.: J/A+AS/120/89 VIII/37 Fich M.: J/AJ/92/787 J/ApJ/559/307 J/ApJS/162/346 FICH M.: VII/50 VII/76 Fichtel C.E.: IX/9 J/ApJS/101/259 J/ApJS/107/227 J/ApJS/123/79 J/ApJS/94/551 Fiedler R.L.: J/ApJS/136/265 Fiegle K.: J/A+AS/98/589 Field D.: J/A+A/453/517 J/A+A/466/949 J/A+A/469/459 Fienga A.: J/A+A/391/767 J/A+A/406/751 Fierro J.M.: J/ApJS/101/259 J/ApJS/107/227 J/ApJS/94/551 Figer D.F.: J/ApJ/581/258 FIGON P.: III/184 Figon P.: III/190B III/213 J/A+AS/131/319

Figueras F.: J/A+A/319/881 J/A+A/326/722 J/A+A/336/953 J/A+A/343/446 J/A+A/425/489

J/A+AS/115/401 Filges L.: J/A+A/327/1177 Filimonov K.: J/other/APh/26.282

Filipovic M.D.: J/A+AS/111/311 J/A+AS/120/77 J/A+AS/121/321 J/A+AS/127/119 J/A+AS/130/421

J/A+AS/136/81 J/A+AS/142/41 J/MNRAS/335/1085 J/MNRAS/355/44 J/other/SerAJ/166.13Short Filippenko A. V.: J/ApJS/107/215 Filippenko A.V.: J/A+A/428/555 FILIPPENKO A.V.: J/AJ/107/1022 Filippenko A.V.: J/AJ/121/3048 J/AJ/129/2138 J/AJ/130/2453 J/AJ/131/1648 J/ApJ/594/1 J/ApJ/600/L99

J/ApJ/602/571 J/ApJ/607/665 J/ApJ/613/200 J/ApJ/627/579 J/ApJ/632/799 J/ApJ/636/1002 J/ApJS/112/315 J/ApJS/115/185 J/ApJS/98/477 J/PASP/110/553 J/PASP/114/820 J/PASP/115/1280 J/PASP/117/773

Filippi M.: J/A+AS/129/583 Finch C.T.: J/AJ/130/1658 Fine S.: J/MNRAS/372/425 Finger M.H.: J/ApJS/154/585 FINK H. H.: J/A+A/274/105 J/A+AS/110/469 Fink H.: J/ApJ/450/51 Fink H.H.: J/A+A/307/33 Finkbeiner D.: J/AJ/121/2308 J/AJ/127/3553 Finkbeiner D.P.: J/AJ/125/1817 J/ApJ/625/6 Finkelstein S.L.: J/ApJ/586/1356 Finley D.S.: J/AJ/107/751 Finn R.A.: J/AJ/121/1840 J/ApJ/604/141 J/ApJ/630/206 Finoguenov A.: J/A+A/419/47 J/A+A/434/801 J/ApJ/622/187 Fiocchi M.: J/ApJ/636/765 J/ApJS/170/175 Fioktistova I.S.: J/A+A/422/925 Fiore F.: J/A+A/362/799 J/A+A/370/900 J/A+A/409/79 J/A+A/432/69 J/A+A/457/501 J/A+A/466/31

J/A+A/469/1211 J/AJ/131/527 J/ApJ/570/100 J/MNRAS/327/771 J/MNRAS/342/575 Fiorenzano A.F.M.: J/MNRAS/363/734

Fiorucci M.: J/A+A/327/61 J/A+A/401/781 J/A+AS/116/403 J/A+AS/117/475 J/A+AS/129/577

J/A+AS/130/109 Firmani C.: J/A+A/288/713 J/A+A/412/633 FISCHEL D.: VI/11 Fischel D.: VI/29A Fischer D.: J/ApJ/637/1102 Fischer D.A.: J/ApJ/622/1102 J/ApJ/634/625 J/ApJS/141/503 J/ApJS/159/141 J/ApJS/169/430 Fischer F.J.: J/A+AS/70/481 Fischer J.-U.: IX/32 FISCHER P.: J/AJ/105/938 J/AJ/106/1508 Fischer P.: J/AJ/110/2166 J/AJ/115/592 J/AJ/116/102 J/ApJ/454/788 J/ApJS/90/83 Fischera J.: J/ApJ/619/755 Fish V.L.: J/ApJS/160/220 Fisher D.: J/AJ/113/950 J/ApJ/459/110 J/ApJ/498/195 J/ApJ/500/714 Fisher J.R.: VII/75 Fisher K.B.: VII/185 VII/185 Fisher R.S.: J/ApJ/636/932 FISHER W.A.: III/79 Fishman G.: J/ApJS/120/399 J/ApJS/156/217 Fishman G.J.: IX/20A J/ApJS/134/385 J/ApJS/154/585 FitzGerald M.P.: II/4 FITZGERALD P.M.: J/AJ/106/1533 Fitzpatrick E.L.: III/195 J/AJ/129/1642 Fitzsimmons A.: J/A+A/444/287 J/A+AS/136/257 J/MNRAS/372/1117 J/MNRAS/379/647 J/MNRAS/379/816

Flaccomio E.: J/A+A/345/521 J/A+A/355/651 J/A+A/455/903 J/A+A/459/199 J/A+A/464/211

J/A+A/474/495 J/ApJ/582/382 J/ApJ/588/1009 J/ApJ/606/466 J/ApJ/608/781 J/ApJS/160/319 J/ApJS/160/450 Flagg L.: J/AJ/127/539 J/AJ/128/3012 Flambaum V.V.: J/MNRAS/345/609 J/other/PASA/19.455 Flanagan J.: J/ApJS/92/53 Flatres T.: J/A+AS/125/399 Flatters C.: J/PASJ/52/997 FLECHTER J.M.: V/40 Fleenor M.C.: J/AJ/130/957 J/AJ/131/1280 Fleming T.A.: J/A+A/316/147 J/A+AS/131/197 J/ApJS/99/701 Flesch E.: J/A+A/427/387 Fletcher A.: J/AZh/73/835 Fletcher A.B.: J/A+A/450/39 FLETCHER J.M.: III/79 Fletcher J.M.: V/64 Fletcher S.: J/AJ/119/1901 J/AJ/127/2436 FLICHE H.H.: VII/32 Flin F.: J/A+AS/146/373 Flin P.: J/A+AS/110/313 J/A+AS/125/459 J/A+AS/94/327 Flint K.P.: J/ApJ/524/536 Fliri J.: J/A+A/445/423 J/A+A/465/375

Floquet M.: III/46 J/A+A/328/551 J/A+A/335/565 J/A+A/409/275 J/A+A/411/565 J/A+A/440/305

J/A+A/451/1053 J/A+A/462/683 J/A+A/472/565 J/A+A/472/577 Florczak M.: J/A+AS/134/463 Floren H.-G.: J/A+A/458/789 Flores H.: J/A+A/382/60 J/A+A/423/867 J/A+A/430/115 J/A+A/435/507 J/A+A/455/107 J/A+A/462/535

J/A+A/462/81 J/A+A/465/1099 J/ApJ/481/49 J/ApJ/517/148 Florio: J/A+A/446/949 Florkowski D.R.: J/AJ/110/880 FLORKOWSKI D.R.: VI/23 Florkowski D.R.: VI/44 FLORSCH A.: V/78 FLORSCH J.: V/78 Flower D.: J/A+A/356/1010 J/A+A/419/999 Flower D.R.: J/A+A/321/293 FLUKS M.A.: J/A+AS/105/311 Flynn C.: J/A+A/317/689 J/A+A/423/517 J/A+A/431/143 FLYNN C.: J/A+AS/109/171 J/A+AS/97/835 Flynn C.: J/AJ/100/1191 J/AJ/117/2329 J/MNRAS/336/879 J/MNRAS/373/13 Focardi P.: J/A+A/310/31 J/A+A/310/8 J/A+A/312/745 J/A+A/335/912 J/A+A/391/35 J/A+A/456/467

J/A+AS/101/259 FOCARDI P.: J/AJ/105/894 Focardi P.: J/AJ/109/173 J/AJ/129/1832 J/MNRAS/364/146 Fodor S.: J/ApJS/155/33 Foellmi C.: J/A+A/463/981 Foerster P.: J/A+A/425/767 Foerster-Schreiber N.M.: J/AJ/125/1107 J/AJ/131/1891 Fogwell S.: J/AJ/130/1977 Fohlman M.: J/A+A/467/665 Fokin A.: J/A+A/379/305 J/MNRAS/362/1167 Folatelli G.: J/A+A/437/789 Foley R.J.: J/AJ/130/2453 J/AJ/131/1648 Folha D.F.M.: J/A+A/365/90 J/A+A/399/271 Folkes S.L.: J/A+A/459/L13 VI/112 Folkmann F.: J/A+A/439/387 Folsom C.P.: J/A+A/475/1053 J/MNRAS/376/361 Foltz C.: J/A+A/419/469 J/ApJ/618/123 Foltz C.B.: J/AJ/109/1498 J/AJ/122/518 J/ApJ/445/62 J/ApJ/450/51 J/ApJS/115/185 J/ApJS/126/133

J/ApJS/134/35 Fomalont E.: J/AJ/129/1163 J/PASJ/52/997 Fomalont E.B.: J/AJ/123/2402 J/AJ/126/2562 J/AJ/131/1872 J/AJ/77/405 FOMALONT E.B.: J/ApJ/405/498 Fomalont E.B.: J/ApJ/616/110 J/ApJS/131/95 J/ApJS/141/13 J/ApJS/155/33 J/MNRAS/358/1159 Fong R.: J/MNRAS/236/207 J/MNRAS/243/1 J/MNRAS/300/417 Fonseca M.V.: J/A+A/406/L9 J/A+A/410/813 Fonseca V.: J/A+A/390/39 Fontaine G.: J/A+A/459/557 FONTAINE G.: J/AJ/105/35 Fontaine G.: J/ApJ/631/1100 Fontana A.: J/A+A/318/347 J/A+A/341/641 J/A+A/375/1 J/A+A/392/395 J/A+A/437/883 J/A+A/449/951

J/A+AS/127/217 J/AJ/120/2206 J/AJ/122/2190 FONTANELLI P.: VII/159 Fontani F.: J/A+A/432/921 J/A+A/447/221 Fonti S.: J/A+A/330/505 Foord M. E.: J/A+A/433/745 Foord M.E.: J/A+A/416/383 J/A+A/424/363 J/A+A/455/1157 Forbes D.: J/AJ/130/1194 Forbes D.A.: J/A+A/455/453 J/AJ/119/727 J/MNRAS/327/1116 J/MNRAS/328/461 J/MNRAS/330/547

J/MNRAS/345/949 J/MNRAS/355/608 J/MNRAS/366/1230 J/MNRAS/369/1351 J/MNRAS/370/1223 Forbrich J.: J/ApJS/161/361 Ford A.: J/AJ/128/1 J/MNRAS/364/272 Ford H. C.: J/A+A/366/498 J/ApJS/122/431 Ford H.: J/A+A/431/793 J/A+A/461/823 J/AJ/110/2739 J/ApJ/463/26 J/ApJ/621/53 J/ApJS/150/1

Ford H.C.: J/AJ/108/1598 J/AJ/122/1330 J/AJ/128/1990 J/AJ/131/143 J/AJ/99/1079 J/ApJ/339/53 J/ApJ/344/704 J/ApJ/344/715 J/ApJ/356/332 J/ApJ/414/463 J/ApJ/449/592 J/ApJ/458/455 J/ApJ/483/745 J/ApJ/507/655 J/ApJ/508/491 J/ApJ/512/48 J/ApJ/514/614 J/ApJ/521/155 J/ApJ/523/540 J/ApJ/602/685 J/ApJ/612/202 J/ApJ/644/30 J/ApJS/150/367 J/ApJS/38/351 J/ApJS/88/423

Ford V.I.: II/34A Ford V.L.: J/ApJS/107/97 J/ApJS/81/413 FORESTINI M.: J/A+A/285/473 Forkert T.: J/A+AS/135/571 VIII/57 Forman W: IX/13 Forman W.: J/ApJ/412/479 J/ApJ/502/558 J/ApJ/519/533 J/ApJ/562/254 J/ApJ/594/154 J/ApJ/635/1053

J/ApJS/111/163 J/MNRAS/292/419 FORMAN W.: VII/18 Forman W.R.: J/ApJ/560/675 J/ApJ/641/140 J/ApJS/161/9 Formanek F.: J/A+A/402/745 Formiggini L.: J/A+A/459/557 J/ApJS/111/143 J/MNRAS/332/441 Forrest D. J.: J/ApJS/120/409 Forrester W.L.: J/AJ/109/1112 Forrester W.T.: II/80 FORSSTROEM V.: J/A+A/288/601 Forster J.R.: J/A+AS/137/43 J/AJ/130/1418 Forster K.: J/AJ/130/825 J/ApJ/619/L95 J/ApJS/134/35 J/ApJS/143/257 J/ApJS/150/165 J/ApJS/164/38 Forster Schreiber N.M.: J/ApJ/604/521 Fort B.: J/A+A/422/407 J/A+A/444/365 FORTE J.C.: J/AJ/105/1762 Forte J.C.: J/AJ/127/2114 J/MNRAS/355/608

Forveille T.: J/A+A/267/515 J/A+A/275/163 J/A+A/331/581 J/A+A/338/1066 J/A+A/364/665 J/A+A/380/590 J/A+A/401/959 J/A+A/441/653 J/A+A/474/293 J/A+A/475/1125 J/A+AS/99/291 J/ApJ/633/630 J/ApJS/160/272

Fosbury R.A.E.: J/A+A/434/53 J/A+A/454/423 J/A+A/478/83 Fossat E.: J/A+A/390/717 Fossati G.: J/A+A/375/739 Fossati L.: J/A+A/445/901 Foster B.E.: III/223 Foster D.C.: J/A+AS/126/81 J/A+AS/143/409 Foster R.S.: J/ApJS/136/265 Foucaud S.: J/A+A/403/857 J/A+A/428/1043 J/A+A/439/413 J/A+A/457/79 Fouque: J/A+A/428/587 FOUQUE P.: II/177 Fouque P.: J/A+A/349/L69 J/A+A/362/215 J/A+A/363/62 J/A+A/371/497 J/A+A/377/428 J/A+A/380/590

J/A+A/383/398 J/A+A/387/1 J/A+A/389/149 J/A+A/401/959 J/A+A/411/381 J/A+A/413/1037 J/A+A/415/521 J/A+A/415/531 J/A+A/454/185 J/A+A/476/73 J/A+AS/102/57 J/A+AS/135/133 J/A+AS/91/371

FOUQUE P.: J/A+AS/99/31 Fouque P.: J/AJ/123/789 J/AJ/128/1167 J/ApJ/552/289 J/ApJS/126/1 VII/119 VII/136 Fouqu/’e P.: VII/137B Fouque P.: VII/155 VII/242 FOURNIER A.: J/A+A/297/494 Fouts G.: III/144 III/145 Fowler J.: II/246 Fox A.J.: J/A+A/473/791 Fox D.C.: J/A+A/310/173 Fox D.W.: J/A+A/399/365 Fox M.: J/MNRAS/344/887 J/MNRAS/351/1290 J/MNRAS/372/1621 Fox Machado L.: J/A+A/468/637 F.R.: J/MNRAS/361/272 FRACASSINI M.: II/155 Fracassini M.: II/61 Fraga L.: J/A+A/457/1033 Fragoulopoulou E.: J/A+A/348/184 Fraley C.: J/ApJ/508/314 Franca F.L.: J/MNRAS/351/1290 J/MNRAS/358/397

Franceschini A.: J/A+A/364/517 J/A+A/372/364 J/A+A/382/60 J/A+A/427/23 J/A+A/427/773 J/A+A/434/149 J/A+A/440/61 J/A+A/457/501 J/A+A/467/565 J/A+A/467/73 J/A+AS/141/257 J/AJ/127/3075 J/AJ/129/1198 J/AJ/129/2074 J/MNRAS/302/222 J/MNRAS/305/297 J/MNRAS/325/1173 J/MNRAS/351/1290 J/MNRAS/355/97 J/MNRAS/358/397 J/MNRAS/371/1891

Franchini M.: J/ApJ/601/485 J/ApJ/613/312 J/ApJ/634/1319 Franciosini E.: J/A+A/446/501 J/A+A/468/353 Francis O.P.: J/ApJS/115/185 Francis P.J.: J/AJ/112/2408 J/AJ/127/646 J/ApJ/614/75 VIII/34 FRANCO G.A.P.: J/A+AS/104/9 J/A+AS/112/95 Franco G.A.P.: J/A+AS/130/133 Franco G.A.P: J/A+AS/93/373 Franco G.A.P.: J/MNRAS/331/474 Franco J.: J/AJ/116/1801 FRANCOIS P.: III/165 Francois P.: III/200B J/A+A/403/1105 J/A+A/416/1117 J/A+A/459/423 J/A+A/464/939 J/A+A/464/953

J/A+A/464/967 Francou G.: J/A+A/329/329 FRANCOU G.: VI/66 FRANCOU G: VI/81 Francou G.: VI/87 VI/88 Frandsen S.: J/A+A/465/965 J/A+AS/140/135 Frankowski A.: J/A+A/464/377 Fransson C.: J/A+A/428/555 J/ApJS/99/223 Franx: J/MNRAS/273/1097

Franx M.: J/AJ/122/2205 J/AJ/125/1107 J/AJ/128/1990 J/AJ/130/1 J/AJ/131/1891 J/AJ/134/1103 J/ApJ/459/110 J/ApJ/498/195 J/ApJ/500/714 J/ApJ/539/577 J/ApJ/604/521 J/ApJ/606/L25 J/ApJ/612/202 J/ApJ/621/53 J/ApJ/624/L5 J/ApJ/644/30 J/ApJ/661/750 J/ApJS/126/271 J/ApJS/126/331 J/ApJS/150/1 J/ApJS/162/1

Franz G.: J/ApJS/160/511 Franz O.G.: J/AJ/111/393 J/AJ/111/936 J/AJ/117/1890 J/AJ/119/3084 J/AJ/120/2638 J/AJ/131/1000

Franzetti P.: J/A+A/361/863 J/A+A/403/857 J/A+A/428/1043 J/A+A/457/79 J/A+AS/142/65

J/A+AS/142/73 J/MNRAS/332/422 Frasca A.: J/A+A/333/205 J/A+A/375/130 J/A+A/388/298 J/A+A/429/309 J/A+A/454/301 J/A+A/459/557

J/A+A/470/281 J/AN/328/938 Fraser O.J.: J/AJ/126/3007 J/AJ/128/463 Frattare L.M.: J/AJ/121/66 J/AJ/123/1292 J/AJ/127/1943 J/AJ/127/1959 Frayer D.: J/AJ/129/2074 J/ApJS/154/155 Frayer D.T.: J/AJ/131/250 Frayn C.: J/MNRAS/330/792 Frazer J.: J/PASP/96/707 Frederick C.L.: II/56 Freedman W.F.: J/ApJ/463/26 Freedman W.L.: J/A+A/415/521 FREEDMAN W.L.: J/AJ/104/1349 Freedman W.L.: J/AJ/106/964 J/AJ/128/2239 J/ApJ/507/655 J/ApJ/508/491 J/ApJ/512/48 J/ApJ/514/614

J/ApJ/521/155 J/ApJ/523/540 J/ApJ/529/723 J/ApJ/571/56 J/ApJS/142/71 J/ApJS/89/85 FREEDMAN W.L.: J/ApJS/91/583 Freeman D.E.: VI/61 Freeman K.: J/AJ/120/2496 J/ApJ/542/257 J/ApJ/598/597 Freeman K.C.: J/A+A/360/472 J/A+A/423/97 J/A+A/445/503 FREEMAN K.C.: J/A+AS/97/835 Freeman K.C.: J/AJ/100/1191 J/AJ/114/326 J/AJ/118/752 J/AJ/123/873 J/AJ/124/1954 J/AJ/124/690

J/AJ/125/2502 J/AJ/127/334 J/AJ/128/16 J/ApJ/449/592 J/ApJ/458/455 J/ApJ/462/241 J/ApJ/515/610 J/ApJ/552/289 J/ApJ/562/337 J/ApJ/602/685 J/ApJS/150/367 J/ApJS/165/307 J/MNRAS/337/641 J/MNRAS/342/738 J/MNRAS/346/787 J/MNRAS/355/1303 J/MNRAS/369/120 V/65

Freeman P.: J/ApJS/150/19 Freeman T.: J/AJ/127/1982 Freire Ferrero R.: J/A+A/454/301 FREIRE FERRERO R.: J/A+AS/108/611 Freire P.C.: J/ApJ/625/796 Fremat Y.: J/A+A/385/986 J/A+A/409/275 J/A+A/411/565 J/A+A/438/201 J/A+A/440/305 J/A+A/451/1053

J/A+A/462/683 J/A+A/472/565 J/A+AS/131/469 J/MNRAS/356/545 Fremy J.R.: J/A+A/468/637 French L.: J/AJ/112/2408 French R.S.: J/AJ/116/2395 Frenk C.S.: J/MNRAS/308/897 J/MNRAS/329/227 J/MNRAS/329/87 J/MNRAS/333/100 J/MNRAS/350/1485

J/MNRAS/352/939 VII/221 Fresneau A.: I/303 I/68A I/96 J/A+A/469/1221 J/AJ/130/2701 FREUDLING W.: J/A+AS/108/621 J/A+AS/112/429 Freudling W.: J/A+AS/140/327 J/AJ/113/1197 J/AJ/113/22 J/AJ/117/1668 J/AJ/117/2039 J/ApJ/449/527 Freund M.: J/A+A/376/997 Frew D.J.: J/MNRAS/362/689 V/127 Frey M.: J/A+A/465/331 Frey S.: J/A+A/342/378 J/ApJS/131/95 J/ApJS/155/33 J/PASJ/52/997 Freyberg M.: J/A+A/434/483 J/A+A/469/27 Freyhammer L.M.: J/A+A/374/1056 Friberg P.: J/ApJS/117/427 Fricke K.J.: J/A+A/316/123 J/A+A/341/709 J/A+A/358/819 J/A+A/398/49 J/A+A/398/89 J/A+A/413/37

J/A+A/420/97 J/A+A/421/41 J/A+A/421/539 J/A+A/429/115 J/A+A/464/885 J/AJ/121/169 Fricke W.: I/143 I/149A I/15 I/175 Fried J.W.: III/175 J/A+AS/98/393 Fried R.: J/A+A/374/1056 Friedemann C.: III/122 III/164 III/95 Friedemann: J/A+AS/120/429

Friedjung M.: J/A+A/258/423 J/A+A/381/507 J/A+A/448/1107 J/A+A/452/503 J/A+AS/136/95

J/A+AS/146/407 Friedlander D.P.: J/ApJS/101/259 J/ApJS/107/227 J/ApJS/123/79 FRIEDMAN H.: VII/80 Friedman P.G.: J/AJ/130/825 J/ApJ/619/L95 J/ApJS/164/38 Friedman S.: J/AJ/123/567 Friedman S.D.: J/AJ/121/2308 J/AJ/121/2331 J/ApJS/146/1 Friedrich S.: J/A+A/363/1040 FRIEL E.: III/165 Friel E.D.: III/200B J/AJ/113/1723 J/AJ/129/2725 J/PASP/109/883 Frieman J.: J/AJ/125/1817 J/AJ/131/2018 Frieman J.A.: J/AJ/121/2308 J/AJ/121/2331 J/AJ/123/567 J/AJ/126/2579 Friesen T.P.: J/A+A/413/285 Frinchaboy P.M.: J/AJ/128/245 J/AJ/130/2677 Fringant A.M.: II/111 II/239 III/64 Frink S.: J/A+A/333/619 J/A+AS/146/323 Fritz A.: J/A+A/421/59 J/MNRAS/337/256 J/MNRAS/358/233 Fritz J.: J/A+A/467/565 Fritz S.: J/A+A/447/245 Fritz T.: J/A+A/332/586 Fritze - von Alvensleben U.: J/A+A/341/709 Fritze-v. Alvensleben U.: J/A+A/316/123 J/A+A/398/89 Fritze-v.Alvensleben U.: J/A+A/358/819 J/A+A/392/1 J/A+A/401/1063 Fritze-V.Alvensleben U.: J/A+A/413/37 F.R.Jr.: J/A+A/341/751 Froebrich D.: J/ApJS/156/169 J/MNRAS/374/399 Froehlich H.-E.: J/A+A/454/295 Froeschle C.: J/A+A/320/631 Froeschle M.: J/A+AS/110/107 J/A+AS/125/399 Froese Fischer C.: J/A+A/385/716 Frogel J.A.: J/AJ/109/1131 J/AJ/109/1154 J/AJ/110/2788 J/AJ/110/2844 J/AJ/131/866 J/ApJ/353/494

J/ApJS/143/73 Frolov M.S.: II/139B II/205 II/214A Frolov V.N.: II/123A FROLOV V.N.: II/185 Frolov V.N.: J/A+A/396/125 J/A+A/401/531 J/A+A/451/901 FROLOV V.N.: V/28 Frolov V.N.: V/52 Fromhold-Treu R.: J/ApJ/473/822 Froning C. S.: J/PASP/111/812 Frontera F.: J/A+A/455/813 J/ApJS/156/217

Fruchter A.S.: J/AJ/112/1335 J/AJ/119/486 J/AJ/120/2747 J/AJ/121/1192 J/MNRAS/323/839

J/MNRAS/359/1524 Fruscione A.: J/AJ/107/751 J/AJ/108/1843 J/ApJS/94/127 Fry A.M.: J/A+A/415/531 J/AJ/113/1073 Fry D.J.I.: J/PASP/106/1120 Frye B.: J/ApJ/621/53 Frye R.L.: J/AJ/77/486 Ftaclas C.: J/AJ/108/1 Fu H.-H.: J/AJ/108/2299 Fu H.W.: J/other/NewA/11.27 Fu J.-N.: J/A+AS/115/469 Fuchs B.: J/AJ/119/2866 Fuente A.: J/A+A/285/247 J/A+A/319/235 Fuerboeck C.: J/A+A/388/704

Fuerst E.: J/A+A/323/739 J/A+A/363/141 J/A+AS/106/303 J/A+AS/109/147 J/A+AS/126/413

J/A+AS/85/805 Fugazza D.: J/A+A/363/29 J/A+A/367/759 J/A+A/385/793 J/A+A/445/901 J/A+A/474/793 Fuhr J.R.: VI/72 VI/72 FUHRMANN K.: J/A+A/300/751 Fuhrmann L.: J/A+A/402/151 Fuhrmeister B.: J/A+A/439/1137 Fuhrmeister: J/A+A/403/247 Fujii T.: J/A+A/385/884 J/PASJ/53/293 J/PASJ/54/719 J/PASJ/56/261 J/PASJ/56/765 J/PASJ/58/529 Fujisawa K.: J/PASJ/52/631 J/PASJ/52/997 Fujishita M.: J/PASJ/52/631 Fujita S.S.: J/PASJ/57/165 Fujiyoshi T.: J/MNRAS/379/663 FUKAYA R.: I/186 Fukaya R.: I/249 Fukazawa Y.: J/PASJ/56/965 Fukuda I.: III/63B

Fukugita M.: J/AJ/121/2308 J/AJ/121/2331 J/AJ/121/31 J/AJ/122/503 J/AJ/123/567 J/AJ/125/1817 J/AJ/126/2209 J/AJ/126/2579 J/AJ/131/49 J/ApJ/588/824 J/ApJS/148/243 J/ApJS/163/80 J/ApJS/97/77 J/PASJ/55/771

Fukui Y.: J/ApJS/110/21 J/ApJS/95/419 J/PASJ/51/791 J/PASJ/53/1081 J/PASJ/53/971 J/PASJ/56/313

J/PASJ/57/917 Fukushi H.: J/MNRAS/364/117 J/MNRAS/370/1979 Fukuzawa K.: J/ApJ/519/697 Fulbright J.P.: J/A+A/461/509 J/AJ/120/1841 J/ApJ/595/1154 J/ApJ/612/447 Fuller G.A.: J/A+A/417/115 J/A+A/440/151 J/A+A/468/1009 J/A+A/472/187 Fullmer M.: II/156A Fullton L.K.: J/AJ/110/652 J/AJ/117/247 Fulton B.: J/MNRAS/367/1609 Funaki T.: J/PASJ/52/631 Funes: J/A+A/416/507 Funes J.G.: J/A+A/360/439 J/A+A/374/394 J/A+AS/131/105 J/ApJ/600/716 Funes J.G.S.J.: J/A+A/342/671 Furesz G.: J/A+A/376/561 J/AJ/129/363 Furia M.: J/A+AS/137/269 Furst E.: J/A+AS/58/197 Furst E: J/A+AS/83/539 Furusawa H.: J/ApJ/634/861 J/MNRAS/357/1357 J/PASJ/55/1079 J/PASJ/57/165 FUSE Science Team: B/fuse Fushiki I.: J/ApJ/416/276 FUSI PECCI F.: J/A+A/290/69

Fusi Pecci F.: J/A+A/298/461 J/A+A/320/757 J/A+A/324/915 J/A+A/413/343 J/A+A/416/917

J/A+A/456/985 J/A+A/471/127 J/A+A/473/429 J/A+AS/102/397 FUSI PECCI F.: J/A+AS/67/327 J/A+AS/84/59 Fusi Pecci F.: J/AJ/109/650 J/AJ/109/663 J/AJ/118/1671 J/AJ/118/2839 FUSI PECCI F.: J/MNRAS/252/357 Fusi Pecci F.: J/MNRAS/264/273 FUSI PECCI F.: J/MNRAS/272/391 FUSI PESCI F.: J/AJ/100/1811

G.: J/A+A/394/395 J/A+A/394/479 J/A+A/395/753 J/A+A/421/937 J/A+A/454/855 J/A+A/469/529

J/A+A/469/529 J/A+AS/136/509 J/A+AS/31/439 J/PASP/110/779 Gaalfalk M.: J/A+A/458/789 Gaba A.E.: J/AJ/123/1216 Gabasch A.: J/A+A/398/49 J/A+A/418/885 J/A+A/420/97 J/A+A/421/41 Gabriel C.: J/A+A/365/409 Gabuzda D.C.: J/A+A/425/99 Gachard M.T.: I/230 Gaedke A.: J/A+A/468/1039 Gaensicke: J/A+A/386/1001 Gaensler B.M.: J/ApJ/626/887 J/ApJ/634/L89 J/ApJS/155/89 J/ApJS/159/242 Gaessler W.: J/ApJS/115/185 Gaetz T.J.: J/ApJS/150/19 Gaffney III J.A.: J/ApJ/574/258 J/ApJ/584/911 Gagen-Torn V.A.: J/AZh/71/717 Gagne M.: J/AJ/121/337 J/ApJ/437/361 J/ApJ/445/280 J/ApJ/598/375 J/ApJ/613/393 J/ApJS/91/625 Gahm G.F.: J/ApJS/165/256 Gaidos E.J.: IX/18 J/ApJ/566/744 Gaisser T.K.: J/other/APh/26.282 Gajdosik M.: J/A+AS/137/293 Gal R.: J/AJ/112/2408 GAL R.R. Astron. J. 105: J/AJ/105/1206 Gal R.R.: J/AJ/125/1660 J/AJ/125/2064 J/AJ/128/1017 J/AJ/130/425

Galadi-Enriquez D.: J/A+A/333/471 J/A+A/337/125 J/A+A/426/819 J/A+A/426/827 J/A+A/437/457

J/A+A/470/585 GALADI-ENRIQUEZ D.: J/A+AS/114/489 Galadi-Enriquez D.: J/A+AS/131/239 J/A+AS/146/169 Galan C.: J/A+A/415/273 Galassi M.: J/other/GCN/1070.1 Galavis M.E.: J/A+AS/123/159 J/A+AS/131/499 J/A+AS/137/165 Galaz G.: J/A+AS/119/413 J/AJ/131/2035 J/ApJS/162/1 J/MNRAS/382/279 Galazutdinov G.: J/PASP/118/636 Galazutdinov G.A.: J/A+AS/136/453 J/ApJ/634/451 J/AZh/81/541 Galazutdinov G.S.: J/A+A/414/949 Galeev A.I.: J/AZh/81/541 Galfalk M.: J/A+A/420/945 J/A+A/475/281 Galis R.: J/A+A/373/950 J/A+AS/135/29 Galisky J.: J/PASP/110/553 Galkin V.D.: III/201 J/PAZh/32/672 Gallagher III J. S.: J/PASP/108/682 Gallagher III J.S.: J/A+A/391/857 J/AJ/117/1700 J/AJ/120/1808 J/AJ/121/768 J/AJ/122/3046 Gallagher Iii J.S.: J/ApJ/603/503 Gallagher J.: J/other/APh/26.282 Gallagher J.S.: J/AJ/123/3154 J/AJ/125/1261 J/MNRAS/327/918 Gallagher J.S.III: J/A+A/445/765 J/A+A/451/1159 J/AJ/113/1001 J/AJ/117/75 J/AJ/129/1465 Galland F.: J/A+A/388/632 Gallart C.: J/AJ/110/212 J/AJ/127/840 J/ApJ/623/181 Gallaway M.: J/A+A/467/785 Gallazzi A.: J/A+A/396/449 J/A+A/446/839 Gallego A.T.: J/A+A/417/615

Gallego J.: J/A+A/365/370 J/A+A/379/798 J/A+AS/118/7 J/A+AS/120/323 J/A+AS/141/409 J/AJ/105/427 J/ApJ/481/673 J/ApJS/105/343 J/ApJS/122/415 J/ApJS/95/387 J/MNRAS/338/508 J/MNRAS/338/525 J/MNRAS/359/930

Galleti S.: J/A+A/416/917 J/A+A/456/985 J/A+A/471/127 J/A+A/473/429 J/AJ/122/2569 Galletta G.: J/A+A/393/389 J/A+A/405/5 J/A+A/420/873 J/A+A/422/941 J/A+AS/124/61 GALLI G.: J/A+AS/108/55 Galli M.: J/A+AS/101/259 Galliani S.: III/216 III/249 Galliano F.: J/A+A/451/57 Gallimore J.F.: J/ApJS/158/139 Gallino R.: J/A+A/417/269 J/ApJ/645/613 Gallozzi S.: J/A+A/449/951 Galue H.A.: V/116 Galvez M.C.: J/A+A/379/976 J/MNRAS/328/45 Gal-Yam A.: J/A+A/402/1157 Gamarova A.Y.: J/MNRAS/353/310 Gamm N.G.: J/AZh/72/447 Gammie C.F.: VIII/10 VIII/28 Gandhi P.: J/A+A/474/473 J/MNRAS/382/279 Gandolfi D.: J/A+A/470/281 Ganesh S.: II/243 J/A+A/349/L69 J/A+A/362/215 J/A+A/400/903 J/A+A/403/141 J/ApJ/552/289 Ganguly R.: J/ApJ/613/129 Ganis M.: J/other/SerAJ/166.13Short Gansicke B.T.: J/MNRAS/373/687 Ganugapati R.: J/other/APh/26.282 Gao H.: J/A+AS/113/51 Gao J.: J/A+A/446/551 Gao Y.F.: J/other/NewA/11.27 J/other/NewA/11.557 Garavini G.: J/A+A/437/789 Garay G.: J/A+A/276/25 J/A+A/426/97 J/MNRAS/363/405 Garca J.A.: J/A+A/478/659 Garcet O.: J/A+A/439/413 J/A+A/467/73 J/A+A/474/473 J/MNRAS/382/279 Garcia A.: J/A+A/378/861 Garcia A.M.: J/A+AS/100/47 GARCIA A.M.: J/A+AS/113/151 Garcia B.: III/148A GARCIA B.: III/149 Garcia B.: III/204 J/MNRAS/348/857 Garcia D.A.: J/MNRAS/342/738 J/MNRAS/346/787 J/MNRAS/355/1303 Garcia de Maria J.M.: J/A+A/304/415 Garcia E.: J/A+A/436/443 J/A+A/449/937 J/A+A/462/507 J/A+A/470/505 J/A+A/472/121 Garcia G.: J/A+A/422/527 Garcia J.: J/AJ/125/1940 J/other/RMxAA/38.129 Garcia J.M.: J/A+A/336/920 J/A+A/340/196 J/A+A/372/588 J/A+A/461/1065 GARCIA J.M.: J/A+AS/101/563 Garcia J.M.: J/A+AS/115/315 Garcia J.M: J/A+AS/122/559 Garcia J.M.: J/A+AS/125/471 J/A+AS/128/251 J/MNRAS/353/310 Garcia J.R.: J/A+AS/109/201 Garcia Lopez R. J.: J/A+AS/129/41 Garcia Lopez R.J.: J/A+A/326/780 J/A+AS/131/431 J/ApJ/527/879 Garcia Lugo G.: J/other/RMxAA/42.75 Garcia M.: J/A+AS/123/1 J/A+AS/123/59 Garcia M.R.: J/AJ/131/527 J/ApJ/577/738 J/ApJ/609/735 J/ApJ/610/247 Garcia R.A.: J/A+A/478/197 GARCIA VARGAS M.L.: J/A+AS/112/35 Garcia-Barreto J.A.: J/AJ/126/1707 J/ApJS/51/67 Garcia-Berro E.: J/A+A/465/249 Garcia-Burillo S.: J/A+A/405/5 Garcia-Dabo C.E.: J/ApJS/122/415 J/MNRAS/359/930 Garcia-Gomez C.: J/A+A/415/849 Garcia-gomez C.: J/A+A/421/595

Garcia-Lario P.: J/A+A/431/779 J/A+A/446/773 J/A+A/458/173 J/A+A/458/533 J/A+AS/126/479

J/A+AS/147/5 GARCIA-LOBO E.: J/A+AS/93/189 Garcia-Melendo E.: II/242 J/A+A/418/249 J/AJ/114/1190 Garcia-Pelayo J.M.: J/A+A/240/262 J/A+AS/71/297 Garcia-Rojas J.: J/ApJS/120/113 J/MNRAS/362/301 J/MNRAS/368/253 Garcia-Ruiz G.: J/other/RMxAA/32.143 Garcia-Varela J.-A.: J/PASP/114/233 Garcia-Vargas L.M.: J/AJ/110/212 Garcia-Vargas M.L.: J/A+A/359/18 J/A+AS/130/513 GARDNER F.F.: VII/50 Gardner J.P.: J/AJ/119/486 J/ApJ/548/127 J/ApJS/146/209 J/ApJS/94/461 J/MNRAS/356/568 Garg A.: J/ApJ/626/L5

Garilli B.: J/A+A/403/857 J/A+A/428/1043 J/A+A/457/79 J/A+A/463/519 J/A+AS/104/271

J/A+AS/137/495 GARILLI B.: J/AJ/107/1629 Garmany C.D.: II/82 J/AJ/103/1205 J/AJ/108/1256 J/AJ/98/1305 Garmire G.: J/AJ/120/1426 J/ApJ/574/258 J/ApJ/584/911 J/ApJS/160/353 J/ApJS/163/306 Garmire G.P.: J/AJ/122/2810 J/AJ/124/2351 J/AJ/126/539 J/AJ/126/632 J/AJ/131/2164 J/ApJ/554/742

J/ApJ/589/225 J/ApJ/613/1179 J/ApJS/160/319 J/ApJS/161/21 Garn T.: J/MNRAS/376/1251 Garnavich P.: J/AJ/131/1648 J/ApJ/594/1 J/ApJ/602/571 Garnavich P.M.: J/A+A/428/555 J/AJ/130/2453 J/AJ/131/527 J/ApJ/613/200 J/ApJ/626/L5 J/ApJ/627/579 Garnett D. R.: J/ApJ/510/104 Garnett D.R.: J/AJ/130/1083 J/ApJ/513/168 J/ApJS/157/371 Garnier R.: J/A+AS/102/57 GARNIER R.: J/A+AS/113/151 Garnier R.: J/A+AS/117/467 J/A+AS/140/89 J/A+AS/144/475 GARNIER R.: J/A+AS/76/101 Garnier R.: J/A+AS/91/371 Garrett M.A.: J/A+A/472/805 Garrid R.: J/A+AS/125/321 Garrido J.L.: J/A+A/452/739

Garrido R.: J/A+A/330/515 J/A+A/336/920 J/A+A/340/196 J/A+A/374/1056 J/A+A/394/97

J/A+A/405/1095 J/A+A/426/247 J/A+A/438/201 J/A+A/472/565 J/A+A/475/1159 GARRIDO R.: J/A+AS/106/21 Garrington S.T.: J/A+A/472/805 J/A+AS/135/273 J/MNRAS/358/1159 GARRISON R.F.: J/AJ/107/1556 Garrison R.F.: J/AJ/126/2048 J/AJ/132/161 J/ApJS/35/111 Garwood R.W.: J/AJ/96/1655 Gary B.: J/other/IBVS/5526 Gary L. Wycoff: I/281 I/301 Garzon F.: II/270 J/A+A/365/110 J/A+A/378/116 J/A+A/379/564 J/AJ/110/212 J/ApJ/491/L31 Gaskell C.M.: J/A+A/371/79 J/ApJS/115/185 Gaspar A.: J/A+A/410/879 Gatewood G.: J/AJ/131/1015 Gatley I.: J/AJ/107/2120 J/AJ/116/1816 Gaudi B.S.: J/ApJ/626/L5 Gaume R.: I/297 Gauss F.S.: I/268 Gaustad J.E.: J/PASP/105/1127 Gautier Iii T.N.: J/ApJS/149/365 Gautier T.N.: J/ApJ/636/1098 J/ApJS/127/1 GAVAZZI G.: J/A+A/283/12 Gavazzi G.: J/A+A/327/522 J/A+A/343/86 J/A+A/343/93 J/A+A/361/863 J/A+A/384/383 J/A+A/396/449

J/A+A/397/421 J/A+A/401/471 J/A+A/402/37 J/A+A/406/867 J/A+A/411/381 J/A+A/417/499 J/A+A/419/47 J/A+A/425/429 J/A+A/430/411 J/A+A/446/839 J/A+A/449/929 J/A+AS/104/271

GAVAZZI G.: J/A+AS/107/285

Gavazzi G.: J/A+AS/112/257 J/A+AS/116/193 J/A+AS/120/489 J/A+AS/120/521 J/A+AS/121/507

J/A+AS/125/293 J/A+AS/137/101 J/A+AS/142/65 J/A+AS/142/73 J/ApJ/634/272 J/MNRAS/332/422 Gavazzi R.: J/A+A/422/407 Gavignaud I.: J/A+A/428/1043 J/A+A/457/79 J/A+A/474/473 Gavilan M.: J/A+A/432/861 Gavras P.: J/A+A/472/1055 Gawiser E.: J/AJ/123/2206 J/AJ/131/2373 J/AJ/134/1103 J/ApJ/621/104 J/ApJS/162/1 Gawronski M.P.: J/A+A/474/1093 Gaylard M.J.: J/A+AS/110/81 Gaynullina E.R.: J/A+A/440/53 Gazeas K.: J/A+A/374/1056 Gaztanaga E.: J/MNRAS/372/1621 Gaztelu A.: J/MNRAS/337/256 Geach J.E.: J/MNRAS/379/1343 Gear W.K.: J/AJ/130/1418 J/MNRAS/350/785 J/MNRAS/372/1621 Geballe T.: J/A+A/384/473 Geballe T.R.: J/AJ/127/3553 J/AJ/131/2722 J/ApJ/637/1067 Gebauer H.J.: J/A+A/390/39 Gebhard M.: J/AJ/127/1158 Gebhardt K.: J/AJ/107/2067 J/AJ/110/1699 J/AJ/113/1026 J/AJ/119/1268 J/AJ/124/3255 J/AJ/127/2114

J/AJ/129/2138 J/AJ/132/447 J/ApJ/454/788 J/ApJ/571/136 J/ApJS/157/175 J/ApJS/159/41 J/MNRAS/355/608 Geckeler R.: J/A+A/373/576 Geenen H.: J/other/APh/26.282 Geers V.C.: J/A+A/411/565

Geffert M.: J/A+A/349/448 J/A+A/357/471 J/A+A/369/511 J/A+A/413/771 J/A+AS/118/277 J/A+AS/119/307 J/A+AS/124/157 J/A+AS/142/373 J/A+AS/144/227 J/A+AS/145/223 J/ApJS/115/185

Geha M.: J/AJ/124/2039 J/AJ/127/334 J/AJ/131/332 J/AJ/131/814 J/ApJ/542/257 J/ApJ/552/289

J/ApJ/562/337 J/ApJ/598/597 Gehrels N.: J/ApJ/633/L77 J/ApJS/120/335 GEHREN T.: J/A+A/300/751 Gehren T.: J/A+A/423/683 Gehrz R.D.: II/95 J/AJ/107/2120 Geier S.: J/A+A/299/39 Geisler D.: J/A+A/396/473 J/A+A/398/479 J/A+A/433/751 J/A+A/442/917 GEISLER D.: J/AJ/105/1762 J/AJ/106/493

Geisler D.: J/AJ/112/2408 J/AJ/116/723 J/AJ/127/2114 J/AJ/127/373 J/AJ/128/712 J/AJ/128/723

J/AJ/130/2677 J/ApJ/559/812 J/MNRAS/358/1215 J/MNRAS/377/300 Gelfand J.D.: J/ApJ/634/L89 J/ApJS/155/89 J/ApJS/159/242 GELLER M. J.: J/ApJ/384/404 Geller M.: J/AJ/105/788 Geller M.J.: J/A+A/312/745 J/AJ/106/1273 J/AJ/109/1458 J/AJ/109/2368 J/AJ/112/1803 J/AJ/113/483

J/AJ/116/1 J/AJ/116/1573 J/AJ/116/2108 J/AJ/118/2561 J/AJ/119/32 J/AJ/119/44 J/AJ/120/2338 J/AJ/122/1289 J/AJ/122/2893 J/AJ/122/714 J/AJ/123/100 J/AJ/123/2976 J/AJ/124/1266 J/AJ/126/1362 J/AJ/126/2152 J/AJ/127/2002 J/AJ/128/1078 J/AJ/130/1097 J/AJ/132/197 GELLER M.J.: J/ApJ/365/66 Geller M.J.: J/ApJ/470/724 J/ApJ/518/69 J/ApJ/534/114 J/ApJ/554/L129 J/ApJ/559/606 J/ApJ/582/668

J/ApJ/607/202 J/ApJ/622/187 J/ApJS/119/277 J/ApJS/121/287 J/ApJS/99/391 J/PASP/111/438 J/PASP/117/227

GELLER M.J.: VII/144 Geller M.J.: VII/164 VII/193 GELLER M.J.: VII/84A VII/86A Gelly B.: J/A+A/390/717 Geminale A.: J/A+A/419/149 GEMMO A.: J/A+AS/112/347 Gendre B.: J/A+A/400/521 J/A+A/455/813 Gennari N.: J/MNRAS/354/708 J/MNRAS/360/185 Genova F.: J/A+A/447/89 Genova R.: J/ApJS/145/355 Genova-Santos R.: J/MNRAS/360/340 Genty V.: III/192 GENTY V.: J/A+AS/110/177 Genty V.: J/A+AS/124/255 J/A+AS/137/451 Genzel R.: J/A+A/443/41 J/ApJ/466/254 J/ApJ/523/248 J/ApJ/552/289 J/ApJ/606/664 J/ApJ/643/1011

J/MNRAS/302/222 J/MNRAS/305/297 J/MNRAS/351/1290 Geoffray H.: J/A+AS/122/415 Geoffrey G. Douglass: I/281 I/301

Georgakakis A.: J/A+A/434/475 J/A+A/459/693 J/A+A/466/823 J/ApJ/624/135 J/ApJ/642/126

J/ApJS/155/1 J/MNRAS/306/708 J/MNRAS/349/135 J/MNRAS/352/91 J/MNRAS/360/782 J/MNRAS/371/221 Georgakakis A.E.: J/MNRAS/366/171 Georgantopoulos I.: J/A+A/434/475 J/A+A/459/693 J/A+A/466/823 J/ApJ/582/615 J/MNRAS/349/135

J/MNRAS/352/91 J/MNRAS/360/782 J/MNRAS/366/171 GEORGELIN Y.: J/A+A/280/365 Georgelin Y.M.: J/A+A/470/161 Georgiev L.: J/AJ/124/3157 Georgiev L.N.: J/A+A/347/583 Ger de Bruyn: VIII/69A Gerard E.: J/A+A/393/1053 J/A+A/414/235 J/A+AS/124/85 Gerasimenko T.P.: V/96 Gerbal D.: J/A+A/310/31 J/A+A/310/8 J/A+AS/111/265 J/A+AS/122/409 GERBALDI M.: II/26A Gerbaldi M.: III/162A J/A+A/379/162 J/A+A/381/105 J/A+A/412/447 J/A+A/472/241 J/A+AS/135/503

J/A+AS/137/273 Gerhard O.: J/MNRAS/369/120 Gerhard O.E.: J/ApJ/635/290 Gerhardt L.: J/other/APh/26.282 Gerin M.: J/A+A/411/381 Gerke B.F.: J/ApJ/625/6 Gerken B.: J/A+A/421/59 Germain M.E.: I/268 I/289 J/AJ/117/1890 J/AJ/117/1905 J/AJ/117/2511 J/AJ/118/1395 J/AJ/119/3071 Germany L.: J/A+A/459/875 Germany L.M.: J/A+A/415/863 J/ApJS/142/71 Gershberg R.E.: J/A+AS/139/555 Gerssen J.: J/AJ/124/3255 Geske M.T.: J/AJ/131/621 J/AJ/131/633 Getman K.: J/ApJS/160/511 Getman K.V.: J/AJ/131/2164 J/ApJ/575/354 J/ApJS/160/319 J/ApJS/160/353 J/ApJS/163/306 Gettel S.J.: J/AJ/131/621 J/AJ/131/633 Geyer E.H.: III/50 Gezari D.: II/98 J/A+AS/134/545 GEZARI D.Y.: II/136 II/184 Gezari D.Y.: II/209 II/216 II/225 GEZARI D.Y.: IV/21 Gherega O.: J/A+AS/125/399 Ghesquiere C.: J/A+A/421/509 Ghez A.M.: J/AJ/123/1570 J/ApJ/636/932 Ghigo F.: J/PASJ/52/997 Ghigo F.D.: II/144 II/145 GHIGO F.D.: J/AJ/102/395 Ghigo F.D.: J/AJ/96/1884 J/ApJS/136/265 J/ApJS/165/439 Ghigo M.: J/MNRAS/367/1170

Ghisellini G.: J/A+A/327/61 J/A+A/375/739 J/A+AS/121/119 J/A+AS/127/445 J/A+AS/129/577

J/A+AS/130/109 Ghosh H.: J/ApJ/618/123 J/ApJS/150/19 Ghosh K.K.: J/ApJS/115/185 J/ApJS/144/213 J/ApJS/154/519 Ghosh S.K.: J/A+A/463/175 J/A+A/468/1001 Ghosh T.: J/A+AS/106/29 J/PASJ/52/997 Giacconi et al.: IX/24 IX/25 IX/26 Giacconi R.: IX/22 IX/3 J/A+A/329/482 J/A+A/371/833 J/A+A/451/457 J/ApJ/595/685 J/ApJS/129/435

J/ApJS/139/369 J/ApJS/155/271 J/ApJS/155/73 J/ApJS/161/21 Giacintucci S.: J/A+A/419/71 GIACONNI R.: VII/18

Giallongo E.: J/A+A/318/347 J/A+A/341/641 J/A+A/392/395 J/A+A/437/883 J/A+A/449/951

J/A+AS/127/217 J/AJ/120/2206 J/AJ/122/2190 J/MNRAS/335/555 Giallongo S.: J/A+A/375/1 Giannakis O.: J/MNRAS/352/91 Gianninas A.: J/ApJ/631/1100 Giannini T.: J/A+A/399/147 J/A+A/419/999 J/A+A/422/141 J/A+AS/136/471 Giannuzzo E.: J/ApJ/510/659 Giardino G.: J/A+A/463/275 Giavalisco A.: J/A+A/454/423

Giavalisco M.: II/261 J/A+A/434/53 J/A+A/478/83 J/AJ/112/1335 J/AJ/127/180 J/ApJ/519/1

J/ApJ/592/728 J/ApJ/600/L99 J/ApJ/607/665 J/ApJ/613/200 Gibb A.G.: J/A+A/325/1163 J/A+A/474/891 J/A+AS/119/333 Gibbons R.: J/A+A/437/789 Gibbs P.: I/144 Gibson B. K.: J/ApJ/514/614

Gibson B.K.: J/AJ/123/873 J/AJ/124/1954 J/AJ/124/690 J/AJ/127/2771 J/AJ/128/16 J/ApJ/507/655 J/ApJ/508/491 J/ApJ/512/48 J/ApJ/521/155 J/ApJ/523/540 J/ApJ/529/723 J/ApJS/142/71 J/ApJS/161/147 J/MNRAS/337/641 J/MNRAS/346/787 J/MNRAS/355/1303

Gibson S.J.: J/ApJ/626/195 GICLAS H.L.: I/112 Giclas H.L.: I/79 Gieles M.: J/A+A/404/L29 Gieren W.: J/A+A/371/497 J/A+A/440/783 J/A+A/445/901 J/A+A/450/873 J/A+A/476/73 J/A+A/478/659

J/AJ/123/789 J/AJ/128/1167 J/AJ/128/2815 J/MNRAS/350/679 J/MNRAS/361/1055 J/PASP/114/298 Gieren W.P.: J/A+A/415/521 J/A+A/415/531 J/A+A/440/487 J/AJ/107/2093 J/ApJS/117/135 Gies D.R.: J/ApJ/525/420 J/ApJ/622/1052 J/ApJ/639/1069 J/ApJS/161/118 Gietzen K.M.: J/A+A/459/557 Gigosos M.A.: J/A+AS/145/491 Gigoux P.: II/179 Gigoyan K.S.: III/237A III/246 J/other/Ap/41.356 J/other/Ap/44.328 J/other/Ap/44.88 J/other/Ap/45.322

J/other/Ap/46.46 J/other/Ap/46.475 Gil de Paz A.: J/A+A/365/370 J/A+AS/141/409 Gil De Paz A.: J/ApJS/147/29 J/ApJS/156/345 J/ApJS/164/38 Gil de Paz A.: J/MNRAS/338/508 Gil De Paz A.: J/MNRAS/338/525 Gilbank D.G.: J/A+A/421/59 J/MNRAS/346/1125 Gilbert A.M.: J/ApJ/581/258 Gilbert K.M.: J/ApJ/613/682 J/ApJS/143/73 Gilfanov M.: J/A+A/418/927 J/A+A/447/71 J/A+A/468/49 J/ApJ/562/254 Gilfanov M.R.: J/PAZh/30/430 Gili R.: J/A+A/369/552 J/A+A/378/954 J/A+A/396/933 J/A+AS/106/377 J/A+AS/126/1 J/A+AS/135/499

J/A+AS/145/67 Gill D.: I/108 GILLESPIE A.R.: VII/50 Gillessen S.: J/ApJ/643/1011 Gillet D.: J/A+A/308/481 J/A+A/379/305 Gillett F.C.: II/126 Gilli G.: J/A+A/449/723 Gilli R.: J/A+A/434/801 J/A+A/451/457 J/ApJS/139/369 J/ApJS/155/271 J/ApJS/155/73 J/ApJS/161/21 Gillieron D.: J/A+A/421/509 Gilliland R.: J/ApJ/612/481 Gilliland R.L.: J/A+A/476/1347 J/AJ/112/1335 J/AJ/128/3053 J/ApJ/559/1060 Gillon M.: J/A+A/466/743 J/A+A/471/L51 J/A+A/474/293 J/A+A/475/1125 Gilmore D.: J/ApJ/581/258 J/ApJ/622/294 J/MNRAS/369/697 Gilmore G.: III/151 III/153 J/A+A/465/357 J/AJ/109/1095 J/AJ/113/634 J/ApJ/552/289 GILMORE G.: J/MNRAS/275/591 Gilmore G.: J/MNRAS/330/792 Gilmore G.F.: J/AJ/131/895 J/MNRAS/345/747

Gilmozzi R.: J/A+A/392/395 J/A+A/437/883 J/A+A/459/875 J/A+A/462/L9 J/ApJS/139/369

J/ApJS/155/271 J/ApJS/99/223 J/MNRAS/327/918 Gim M.: J/PASP/110/1172 Gimenez A.: J/A+A/326/1044 J/A+A/402/509 J/A+A/461/1065 GIMENEZ A.: J/A+AS/101/563 J/A+AS/110/329 J/A+AS/112/123 Gimenez A.: J/A+AS/114/247 J/A+AS/115/315 J/A+AS/122/559 J/A+AS/125/471 J/A+AS/128/251 VI/118 Gimeno G.N.: J/AJ/128/62 Gimeno J.: J/PASJ/52/997 Gimnez A.: J/A+AS/109/425 Ginestet N.: III/217 GINESTET N.: IV/16 Ginestet N.: J/A+A/354/881 J/A+AS/123/135 J/A+AS/130/223 J/ApJS/143/513 GINESTET N.: V/49 Ginestet N.: V/60 Gioa I.M.: IX/13 Gioia I.: J/A+A/461/977

Gioia I.M.: IX/15 J/A+A/446/97 J/A+A/462/449 J/ApJ/594/154 J/ApJ/617/192 J/ApJS/149/29

J/ApJS/162/304 Giommi P.: IX/12 IX/31 J/A+A/362/799 J/A+A/434/385 J/A+A/448/L9 J/A+A/470/787 J/A+A/472/699

J/A+A/472/705 J/A+A/473/347 J/A+AS/109/267 J/AJ/115/1253 J/ApJ/570/100 J/MNRAS/277/1477 J/MNRAS/323/757 J/MNRAS/327/771 J/MNRAS/351/83

Giordano L.: J/ApJ/634/272 Giorgi E.: J/other/NewA/12.461 Giorgi E.E.: J/A+A/381/884 J/A+A/417/961 J/A+A/432/491 J/other/RMxAA/39.89 Giovanardi C.: J/A+A/414/905 J/A+AS/103/541 J/A+AS/137/101 J/AJ/113/905 Giovanelli R.: J/AJ/105/1251 J/AJ/105/1271 J/AJ/113/1197 J/AJ/113/22 J/AJ/113/905 J/AJ/114/122

J/AJ/116/2728 J/AJ/116/2738 J/AJ/117/1668 J/AJ/117/2039 J/AJ/118/1468 J/AJ/118/1489 J/AJ/124/100 J/AJ/127/3273 J/AJ/129/2119 J/AJ/130/1037 J/AJ/130/2613 VIII/77

Giovannelli F.: J/A+AS/114/341 J/A+AS/136/95 Giovannetti P.: J/A+A/330/990 GIOVANNINI G.: J/A+A/281/375 Giovannini G.: J/A+A/370/409 J/A+AS/74/315 Girard P.: J/A+A/438/139 J/A+A/463/265 Girard T.: J/AJ/127/1158 GIRARD T.M.: I/212 Girard T.M.: I/277 I/288 J/AJ/117/548 J/AJ/122/1486 J/AJ/124/601 J/AJ/126/2922 J/AJ/131/1000 Girardi L.: J/A+A/361/1011 J/A+A/361/1023 J/A+A/392/741 J/A+AS/141/371 Girardi M.: J/A+A/397/431 J/A+A/416/839 J/A+A/424/79 J/A+A/432/381 J/A+A/449/461 J/A+A/455/45

J/A+A/467/37 J/A+A/469/861 J/ApJ/569/720 Giraud E.: J/A+A/436/457 Giraud-Heraud Y.: J/A+A/402/113 J/A+A/421/509 Giraud-heraud Y.: J/MNRAS/351/1071 Giridhar S.: J/A+A/368/250 J/A+A/408/L29 J/ApJ/479/427 Girish V.: J/A+A/455/303 Girma M.: J/A+A/406/L9 J/A+A/410/813 Giro E.: J/A+A/434/1107 Girouard F.R.: J/AJ/107/751 Gisle R G.: J/AJ/119/1901 Gisler G.: J/AJ/127/2436 J/ApJ/252/102 Gispert R.: J/A+A/372/364 Giuricin G.: J/A+A/310/31 J/A+A/310/8 J/ApJ/543/178 J/ApJ/569/720 Giveon U.: J/A+A/369/57 J/AJ/129/348 J/AJ/130/156 J/ApJ/566/880 Gizis J.: II/246

Gizis J.E.: III/198 J/AJ/113/806 J/AJ/123/3356 J/AJ/125/3302 J/AJ/130/1871 J/ApJ/564/421

J/PASP/109/849 J/PASP/117/676 J/PASP/118/659 Gladders M.D.: J/ApJ/582/1123 J/ApJS/157/1 J/ApJS/158/161 Gladman B.: J/ApJ/627/621 Glagolevskij Yu.V.: J/A+A/324/949 J/other/BSAO/38.152

Glass I.S.: J/A+A/362/215 J/A+A/453/535 J/ApJ/552/289 J/MNRAS/321/77 J/MNRAS/337/519

J/PASJ/56/765 Glass S.: J/PASJ/56/261 Glauser A.: J/A+A/468/353

Glazebrook K.: II/211 J/AJ/116/102 J/AJ/127/2455 J/AJ/127/3553 J/AJ/131/2383 J/ApJ/499/112 J/ApJ/549/820 J/ApJS/107/1 J/MNRAS/273/157 J/MNRAS/297/885 J/MNRAS/328/150 J/MNRAS/329/227 J/MNRAS/329/87 J/MNRAS/333/100 J/MNRAS/335/10 J/MNRAS/350/1485 J/MNRAS/352/939

Glazebrook K.G.: J/AJ/112/359 GLEBOCKI R.: III/120 Glebocki R.: III/226 III/244 Gleissner T.: J/A+A/407/1039 Glenn J.: J/ApJ/638/293 Glicenstein J-F.: J/A+A/389/149 Glicenstein J.F.: J/A+A/454/185 GLIESE W.: V/1 V/35 V/35 V/70A Glikman E.: J/ApJ/607/60 Glindemann A.: J/A+A/314/846 J/A+A/325/159 Gliozzi M.: J/A+A/433/1163 Glowienka L.: J/A+A/465/965 Glushkova E.A.: J/A+AS/139/47 Glushkova E.V.: III/229 J/A+A/399/105 J/PAZh/23/674 J/PAZh/23/90 J/PAZh/23/931 J/PAZh/25/115

J/PAZh/25/689 Glushkova I.A.: J/AZh/84/227 Glushneva I.N. et al.: III/207 Glushneva I.N.: III/208 III/209 J/A+AS/92/1 Gnacinski P.: III/226 III/244 Gnevysheva K.G.: I/44 Goad M.R.: J/A+A/448/L9 Gochermann J.: J/A+AS/121/247 J/A+AS/99/591 J/MNRAS/306/279 GODET L.: I/74 Godet O.: J/A+A/448/L9 Godfrey P.D.: J/A+A/459/317 Godwin J.G.: J/MNRAS/239/459 J/MNRAS/267/431 VII/42A Goebel F.: J/A+A/410/813 Goecking K.: II/202 Goecking K.-D.: J/A+A/304/415 Goetting N.: J/A+A/390/39 J/A+A/406/L9 J/A+A/410/813 Goetz D.: J/A+A/458/21 Golay M.: J/A+A/268/419 J/other/Ap+SS/163.275 Goldader J.D.: J/ApJS/73/661 GOLDES G.: J/A+A/280/365 Goldhaber G.: J/A+A/437/789 Goldin A.: J/ApJ/638/293 Goldman B.: J/A+A/389/149 J/A+A/454/185 Goldoni P.: J/A+A/458/21 Goldschmidt A.: J/other/APh/26.282 GOLDSCHMIDT P.: J/A+AS/112/347 Goldschmidt P.: J/ApJ/511/612 J/MNRAS/351/1290 Goldstein G.: J/MNRAS/335/1085 Goldsten J.: J/ApJS/156/217 Goldwurm A.: J/A+A/458/21 Golenetskii S.: J/ApJS/156/217 Golev V.: J/A+AS/116/447 J/A+AS/132/255 J/A+AS/136/519

Golimowski D.A.: J/AJ/121/2189 J/AJ/123/3409 J/AJ/127/3553 J/AJ/128/1990 J/AJ/131/2722

J/ApJ/612/202 J/ApJ/621/53 J/ApJ/637/1067 J/ApJS/150/1 Gol’nev V.Ya.: J/AZh/62/229 Gol’neva N.E.: VIII/33 Golombek D.: J/ApJS/122/81 J/ApJS/158/139 Golub L.: J/ApJ/315/687 Golubiatnikov G.Yu.: J/PAZh/33/143 Golubinskiy Y.V.: J/ApJ/622/129 Golubinsky Y.V.: J/A+A/437/87 J/other/Ap/48.304 Golwala S.: J/ApJ/638/293 Gombert G.: II/230 Gomez A.: I/196 J/A+A/331/81 J/A+A/465/799

Gomez A.E.: J/A+A/319/881 J/A+A/326/722 J/A+A/331/949 J/A+A/336/953 J/A+A/381/105

J/A+A/393/897 J/A+A/463/671 J/A+AS/135/503 J/A+AS/137/273 J/A+AS/137/451 Gomez de Castro A.I.: J/A+AS/137/351 Gomez: J/A+A/440/487 Gomez J.F.: J/ApJ/628/789 Gomez J.L.: J/AJ/130/1418 Gomez M.: J/A+A/376/907 J/A+A/399/995 J/AJ/121/2673 J/AJ/121/974 J/AJ/131/527 J/ApJS/117/135

J/MNRAS/367/513 J/PASP/108/332 J/PASP/108/338 Gomez P.: J/AJ/126/2209 Gomez P.L.: J/AJ/120/2269 Gomez X.: J/AJ/112/2408 Gomez-Forrellad J.M.: J/A+AS/125/399 Gomez-Gonzalez J.: J/AZh/71/37 Gonano-Beurer M.: J/A+AS/113/115 Goncalves A.C.: J/A+AS/135/437 J/other/Ap/42.1 Goncalves R.: J/A+A/465/331 Goncalves T.S.: J/AJ/130/425 Goncharov G.A.: J/AZh/83/821 Gondoin P.: J/A+A/438/291 J/A+A/454/595 Gondolo P.: J/A+A/421/509 Gonidakis I.: J/A+A/451/431 J/A+A/476/L17 Gonnella A.M.: J/AJ/119/486 Gontcharov A.G.: I/242 Gontcharov G.A.: I/266 III/252 Gontier A.-M.: J/A+A/375/661 J/AJ/127/3587 Gonzales L.E.: J/AJ/112/2408 Gonzalez A.H.: J/ApJ/563/629 J/ApJ/566/103 J/ApJ/639/816 J/ApJS/137/117 Gonzalez Delgado R.M.: J/A+A/416/475 J/A+A/441/981

Gonzalez G.: II/131 J/A+A/336/263 J/AJ/114/341 J/AJ/118/1245 J/AJ/127/373 J/ApJ/479/427

J/MNRAS/378/1141 J/PASP/111/335 Gonzalez J.: J/A+A/390/39 Gonzalez: J/A+AS/139/29 Gonzalez J.F.: J/PASP/108/338 Gonzalez J.J.: J/AJ/130/55 J/ApJS/116/1 GONZALEZ J.J.: J/ApJS/94/687 Gonzalez L.E.: J/ApJS/133/119 Gonzalez M.A.: J/A+AS/145/491 Gonzalez R.A.: J/ApJ/611/270 Gonzalez V.: J/ApJ/621/104 Gonzalez-Alfonso E.: J/A+A/329/219 Gonzalez-Bedolla S.F.: J/A+A/328/235 Gonzalez-Delgado R.M.: J/AJ/110/212 J/MNRAS/357/945 Gonzalez-Fernandez C.: II/270 Gonzalez-Lopezlira R.A.: J/AJ/119/486 J/AJ/120/2747 GONZALEZ-PATINO D.: J/A+AS/104/557 Gonzalez-Perez J.N.: J/AJ/122/2055 Gonzalez-Riestra R.: J/A+A/373/730 J/A+A/420/571 J/ApJS/95/387 Gonzalez-Serrano I.: J/A+A/339/34 Gonzalez-Serrano J.I.: J/A+AS/142/353 J/MNRAS/311/456 Gonzalez-serrano J.I.: J/MNRAS/348/857 Gonzalez-Serrano J.I.: J/MNRAS/351/1290 Gonzalez-serrano J.I.: J/MNRAS/353/211 Gonzalez-Serrano J.I.: J/MNRAS/368/1411 J/MNRAS/370/1034 Gonzalez-Solares E.: J/AJ/129/2074 J/MNRAS/343/293 J/MNRAS/355/97 J/MNRAS/358/397 J/MNRAS/379/1343 Gonzalez-Solares E.A.: J/AJ/129/1198 J/MNRAS/351/1290 J/MNRAS/358/333 Gonzalez-Solarez E.: J/A+A/472/805 Goobar A.: J/A+A/437/789 J/ApJ/561/106 Goodman A.A.: J/ApJ/517/264 Goodman J.: J/AJ/127/899 Goodman J.A.: J/other/APh/26.282 Goodrich R.W.: J/AJ/127/3121 GOODS Team: J/A+A/454/423 J/A+A/478/83 GOOGE W.D.: I/73 Googe W.D.: I/90 Goorvitch D.: J/ApJS/95/535 Goossens K.: J/A+A/463/243 Gopal Krishna: J/A+A/434/887 Gopal-Krishna: J/A+A/436/457 J/A+AS/106/29 Gopka V.: J/A+A/430/255 J/MNRAS/359/865 Goranskii V.P.: J/AZh/79/309 J/AZh/84/147 J/PAZh/28/201 J/PAZh/28/764 Goranskij V.P.: II/139B II/140 II/205 Goransky V.P.: J/AZh/78/408 J/AZh/79/1087 J/PAZh/26/27 Gorbaneva T.I.: J/AZh/84/429 J/AZh/86/158 Gordon D.: J/AJ/126/2562 J/AJ/129/1163 J/AJ/131/1872 J/ApJS/141/13 Gordon K.D.: J/AJ/131/250 J/ApJ/633/871 J/ApJ/636/1098 J/ApJS/143/377 J/ApJS/154/155 Gordon S.M.: J/ApJS/120/247 Gore G.D.: II/68 Gorel L.: I/308 Gorel L.F.: I/59 Gorgas J.: J/A+A/474/763 J/A+AS/139/29 J/MNRAS/371/703 J/MNRAS/374/664 Gorgas: J/MNRAS/326/959 Goriely S.: J/A+A/342/881 J/A+A/430/255 J/A+A/441/1195 Gorlova N.: J/ApJ/620/1010 Gorlova N.I.: J/A+A/411/559 GORNY S.K.: J/A+A/303/893 Gorny S.K.: J/A+A/318/256 J/A+A/327/736 J/A+A/405/627 J/A+A/427/231 Gorosabel J.: J/A+A/427/87 J/A+A/474/793 J/ApJ/641/L13 Gorshanov D.L.: J/A+A/396/125 J/A+A/451/901 Gorshkov A.G.: J/AZh/77/407 J/AZh/80/978 J/AZh/82/420 J/AZh/83/241 VIII/49 Gorshkov: J/AZh/72/291 Gorti U.: J/PASP/118/1690 Gorynya N.A.: II/139B II/140 II/205 III/229 J/A+A/409/245 J/PAZh/24/443 Gorynya N.G.: J/A+A/374/227 J/A+A/465/257 Goss V.M.: J/AZh/72/613 Goss W.M.: J/AJ/104/704 J/AJ/121/1192 J/AJ/127/3399 J/ApJS/107/239 J/ApJS/120/247 J/ApJS/155/577

J/other/AuJPA/14.1 J/other/AuJPA/14.133 J/other/AuJPA/14.77 J/other/BSAO/40.5 VII/191 Gosset E.: J/A+A/395/499 J/A+A/418/841 J/A+A/439/413 J/A+A/454/1047 J/A+A/463/981

J/A+A/474/473 J/A+AS/123/529 J/A+AS/90/451 J/MNRAS/372/661 J/MNRAS/382/279 Goto M.: J/AJ/125/1537 J/ApJ/566/974 Goto T.: J/AJ/123/1807 J/AJ/128/1990 J/AJ/130/968 J/ApJ/621/53 J/ApJS/148/243 J/MNRAS/360/587

J/PASJ/55/771 GOTTLIEB D.M.: V/47 Gottloeber S.: J/A+A/440/53 Gottwald M.: J/A+AS/109/9 Gotz D.: J/ApJS/170/175 GOUDFROOIJ P.: J/A+A/279/61 Goudfrooij P.: J/A+A/415/123 J/A+A/474/763 GOUDFROOIJ P.: J/A+AS/104/179 Goudfrooij P.: J/A+AS/96/389 J/MNRAS/343/665 J/MNRAS/369/697 Goudis C.: J/ApJ/639/37 Gouguenheim L.: J/A+A/113/61 J/A+A/430/373 J/A+AS/102/57 J/A+AS/130/333 J/A+AS/91/371 VII/119

VII/136 Gouguenheim: VII/238 GOUIFFES CH.: J/A+AS/112/347 Gould A.: J/A+A/389/149 J/A+A/454/185 J/AJ/126/472 J/AJ/130/1916 J/ApJ/409/28 J/ApJ/582/1011

J/ApJS/150/455 J/ApJS/152/103 J/MNRAS/351/1071 Gould D.M.: J/MNRAS/273/411 J/MNRAS/301/235 Gouliermis D.: J/A+A/405/111 Gouliermis D.A.: J/ApJS/166/549 Goupil M.J.: J/A+A/405/1095 GOURGOULHON E.: J/A+A/291/155 J/A+AS/108/455 Gouttebroze P.: J/A+A/313/949 Governato F.: J/MNRAS/351/265 Gower A.C.: J/ApJ/427/125 GOWER J.F.R.: VII/53 VIII/4 GOY G.: II/76A Graczewski K.: J/A+A/447/921 Grado A.: J/A+A/428/339 Grady C.: J/ApJ/612/481 Grady C.A.: J/A+A/365/110 J/A+A/378/116 J/A+A/379/564 Graff D.: J/A+A/389/149 J/A+A/454/185 Graham A.W.: J/A+AS/133/325 J/ApJ/602/664 Graham D.A.: J/A+A/476/L17 GRAHAM J.A.: III/7 Graham J.A.: J/ApJ/463/26 J/ApJ/507/655 J/ApJ/508/491 J/ApJ/512/48 J/ApJ/514/614 J/ApJ/521/155

J/ApJ/523/540 J/ApJ/529/723 J/ApJS/142/71 J/PASP/86/829 J/PASP/94/244 Graham: J/A+A/364/391 Graham J.R.: J/ApJ/475/445 J/ApJ/581/258 Graham K.: J/other/IBVS/5526 Graham M.: J/MNRAS/351/1290 Graham M.J.: J/MNRAS/309/48 Graham P.W.: J/AJ/121/2159 Grainge K.: J/MNRAS/341/1066 J/MNRAS/360/340 Grainge K.J.B.: J/MNRAS/342/915 Granata S.: J/A+A/472/705 Granato G.L.: J/A+A/364/517 J/MNRAS/351/1290 J/MNRAS/372/1621 Grankin K.: J/A+A/463/1017 GRANKIN K.N.: J/A+A/282/L9 Grankin K.N.: J/A+A/350/513 J/A+A/374/1049 J/A+A/409/169 J/A+A/461/183 GRANKIN K.N.: J/A+AS/103/183 J/A+AS/107/9 Grankin K.N.: J/A+AS/139/513 J/AZh/78/408 J/PAZh/26/27 J/PAZh/27/83 J/PAZh/27/942 Granot J.: J/ApJ/634/L89 J/ApJ/641/L13 Grant J.K.: J/AJ/130/2043 Grant N.I.: J/MNRAS/363/1019 Granzer T.: J/A+AS/140/29 J/A+AS/142/275 Grasdalen G.L.: II/95 Grashuis R.: J/AJ/112/742 Gratton: J/A+A/421/937 GRATTON L. (Frascati): VI/26 Gratton R.: J/A+A/430/603 J/A+AS/82/71 J/AJ/110/2319 GRATTON R.G.: J/A+A/283/911 J/A+A/287/927

Gratton R.G.: J/A+A/314/191 J/A+A/350/955 J/A+A/354/169 J/A+A/356/238 J/A+A/377/123 J/A+A/384/879 J/A+A/404/187 J/A+A/413/343 J/A+A/420/683 J/A+A/434/235 J/A+A/440/901 J/A+A/454/581 J/A+A/455/271 J/A+A/464/927 J/A+A/464/939 J/A+A/464/953 J/A+A/464/967 J/A+AS/121/95 J/AJ/121/327 J/MNRAS/363/734

Grauer A.D.: J/ApJS/149/327 Gravel P.: J/ApJS/102/289 J/ApJS/116/231 Graves G.J.M.: J/AJ/131/527 Gray A.G.: J/ApJS/155/257 Gray J.: J/AJ/126/2579 Gray M.: II/253 Gray P.M.: J/ApJS/72/715 Gray R.O.: J/A+AS/147/5 GRAY R.O.: J/AJ/107/1556 Gray R.O.: J/AJ/117/2308 J/AJ/121/2148 J/AJ/121/2159 J/AJ/126/2048 J/AJ/132/161 Grazian A.: J/A+A/449/951 J/A+A/457/501 J/A+A/478/83 J/AJ/119/2540 J/AJ/122/2190 J/AJ/124/2955 Grazian M.: J/A+A/454/423 Graziani C.: J/other/GCN/1070.1 Graziani M.: J/A+A/283/319 Greason M.R.: J/AJ/113/1011 Greaves J.: J/MNRAS/355/585 Greaves S.C.: J/A+AS/121/65

Grebel E.K.: J/A+A/317/448 J/A+A/391/547 J/A+A/396/473 J/A+A/398/479 J/A+A/433/751

J/A+A/445/765 GREBEL E.K.: J/A+AS/109/293 Grebel E.K.: J/A+AS/109/313 GREBEL E.K.: J/A+AS/112/367 Grebel E.K.: J/A+AS/127/167 J/AJ/122/248 J/AJ/123/855 J/AJ/125/197 J/AJ/127/1588 J/AJ/127/1811

J/AJ/127/2114 J/AJ/127/704 J/AJ/128/1606 J/AJ/131/895 J/ApJ/588/824 J/ApJ/598/260 J/ApJ/633/844 J/ApJS/153/429

Grebel R.K.: J/A+AS/135/391 Grebenev S.A.: J/PAZh/30/430 J/PAZh/32/508 Grec G.: J/A+A/390/717 Gredel R.: J/A+A/276/25 J/A+A/372/95 J/A+A/426/171 J/A+A/447/609 J/A+A/474/941 Green A.: VIII/82 Green A.J.: J/A+AS/118/329 J/AJ/124/1954 J/AJ/128/16 J/ApJ/626/887 J/ApJS/155/595 GREEN D.A.: J/A+AS/88/355 Green D.A.: J/MNRAS/274/324 J/MNRAS/294/607 J/MNRAS/301/203 J/MNRAS/376/1251 VII/133 VII/133

VII/163 VII/187 VII/210 VII/211 VII/227 GREEN D.A.: VIII/24 Green E.M.: J/A+A/459/557 GREEN E.M.: VI/40 Green J.C.: J/ApJS/146/1 Green P.: J/ApJS/161/9 Green P.J.: J/A+A/375/366 J/ApJ/450/51 J/ApJ/618/123 J/ApJ/641/140 J/ApJ/644/829 J/ApJ/645/955

J/ApJS/134/35 J/ApJS/143/257 J/ApJS/150/165 J/ApJS/150/19 Green R.: J/AJ/129/2138

Green R.F.: II/207 J/A+A/374/907 J/A+AS/142/417 J/AJ/108/1163 J/AJ/114/2269 J/AJ/123/578 J/AJ/130/873 J/ApJ/269/352 J/ApJ/462/614 J/ApJS/104/185 J/ApJS/119/189 J/ApJS/121/417 J/ApJS/146/1 J/ApJS/95/1

GREEN S.: VI/51 VI/56 VI/59 Greenberg J.M.: J/A+A/275/549 Greene J.E.: J/ApJ/641/L21 Greene M.G.: J/other/APh/26.282 Greene T.P.: J/AJ/127/1131 J/AJ/130/1145 J/AJ/130/1733 Greenfield P.: J/A+A/287/769 J/A+A/288/413 Greenfield P.E.: J/ApJS/51/67 Greenhill L.J.: J/ApJ/581/325 J/ApJ/615/702 Greenlees E.: J/ApJ/612/511 Greenwich: I/34 Gregg M.: J/A+A/451/1159 J/A+A/459/679 Gregg M.D.: J/A+A/356/445 Gregg M.D: J/A+A/476/59 Gregg M.D.: J/AJ/112/407 J/AJ/121/1872 J/AJ/131/1 J/ApJ/607/60 J/ApJS/126/133 J/ApJS/135/227

J/MNRAS/344/188 VIII/48 VIII/51 VIII/59 VIII/71 GREGGIO L.: J/AJ/105/894 Greggio L.: J/AJ/109/173 GREGGIO L.: J/MNRAS/272/391 Gregorich D.: II/156A

Gregorini L.: J/A+A/369/380 J/A+A/379/755 J/A+A/403/857 J/A+A/406/579 J/A+A/428/1043 J/A+A/435/863 J/A+A/457/517 J/A+A/457/79 J/A+A/463/519 J/A+AS/101/259 J/A+AS/106/1 J/A+AS/120/89 J/A+AS/133/129 J/A+AS/139/359 J/A+AS/95/1

GREGORINI L.: J/AJ/105/853 Gregorini L.: VIII/63 GREGORIO-HETEM J.: J/A+A/297/503 Gregorio-Hetem J.: J/A+A/364/674 Gregory P.C.: J/ApJS/90/173 VIII/14 VIII/40 Gregory S.A.: J/AJ/109/2376 Greimel R.: J/A+A/466/1031

Greiner J.: J/A+A/310/384 J/A+A/316/147 J/A+A/361/1005 J/A+A/389/439 J/A+A/397/711

J/A+A/465/375 J/ApJ/610/247 J/other/NewA/5.137 Greisen E.W.: VIII/65 Grenier I.A.: VIII/39

Grenier S.: J/A+A/319/881 J/A+A/326/722 J/A+A/336/953 J/A+A/393/897 J/A+AS/135/503

J/A+AS/137/273 J/A+AS/137/451 Grenon M.: I/196 Grether D.: J/ApJ/640/1051 GREVE A.: J/A+A/293/703 Greve A.: J/A+AS/102/611 J/A+AS/122/271 Greve: J/A+A/364/391 Greve T.R.: J/MNRAS/372/1621 Grevesse N.: J/A+A/347/348 Grewing: J/A+A/364/391 VI/107 Griessmeier J.-M.: J/A+A/475/359 Griest K.: J/AJ/114/326 J/AJ/127/334 J/ApJ/542/257 J/ApJ/552/289 J/ApJ/562/337 J/ApJ/598/597 Griffin K.E.: J/AJ/103/267 Griffin R.: III/169A GRIFFIN R.: J/A+AS/42/391 J/A+AS/42/391 Griffin R.E.: III/169A Griffin R.F.: II/41 J/ApJS/147/103 J/MNRAS/371/1140 J/MNRAS/371/1159 V/103 Griffith M.: J/ApJS/75/801 VIII/52B GRIFFITH M.R.: VIII/27 VIII/27 VIII/27 Griffith M.R.: VIII/38 VIII/38 VIII/38 VIII/38 Griffiths R. E.: J/PASP/108/682 Griffiths R.: J/AJ/113/1001 J/AJ/116/102 J/MNRAS/297/885 Griffiths R.E.: J/AJ/117/1700 J/AJ/122/2810 J/ApJ/453/48 J/ApJ/554/742 J/ApJS/107/1 J/ApJS/121/417

VII/19 Griffiths W.K.: J/A+AS/145/365 GRIFFITHS W.K.: J/MNRAS/250/314 Griffiths W.K.: J/MNRAS/250/683 J/MNRAS/299/1 Grigahcene A.: J/A+A/450/715 J/A+A/456/261 J/A+A/465/965 Grillmair C.: J/AJ/113/1001 Grillmair C.J.: J/AJ/117/1700 J/AJ/119/727 J/AJ/129/2138 J/ApJ/478/124 Grimes J.P.: J/ApJ/475/469 J/ApJS/150/19 Grimm H.-J.: J/ApJS/161/271 Grindlay J.E.: J/ApJ/625/796 J/ApJS/154/585 J/other/Sci/292.2290 Grinin: J/AZh/77/420 Grinin V.: J/A+A/448/1075 Grinin V.P.: J/A+A/327/145 GRININ V.P.: J/A+AS/112/457 Grinin V.P.: J/AZh/73/194 J/AZh/78/60 Grison P.: J/A+A/303/137 GRISON P.: J/A+AS/109/447 Groenewegen M.A.T: J/A+A/317/503 Groenewegen M.A.T.: J/A+A/357/225 J/A+A/379/436 J/A+A/379/740 J/A+A/384/81 J/A+A/390/501

J/A+A/390/511 J/A+A/392/741 J/A+A/406/51 J/A+A/425/595 J/A+A/439/559 J/A+A/443/143 J/A+A/448/181 J/A+A/463/579 J/A+AS/101/513 J/A+AS/123/305 J/A+AS/125/419 J/MNRAS/369/751 J/MNRAS/376/313

Grogin N.: J/ApJ/613/200 J/ApJS/139/369 Grogin N.A: J/AJ/118/2561 Grogin N.A.: J/AJ/119/32 J/AJ/131/527 J/ApJ/595/685 J/ApJ/640/603 J/ApJS/119/277 Groh J.H.: J/A+A/465/993 Gromadzki M.: J/A+A/463/703 Gronbech B.: II/30 II/33 Groner P.: J/ApJ/500/1059 Groner T.: J/AJ/131/527 Gronwall C.: J/AJ/121/66 J/AJ/123/1292 J/AJ/127/1943 J/AJ/127/1959 J/AJ/128/1990 J/AJ/130/2571

J/AJ/130/2584 J/AJ/130/496 J/ApJ/481/673 J/ApJ/612/202 J/ApJ/621/53 J/ApJS/150/1 J/ApJS/159/41 Groom D.E.: J/A+A/437/789 Groot P.J.: J/A+A/466/1031 J/A+AS/118/545 J/AJ/131/527 Gros M.: J/A+A/303/137 J/A+A/389/149 J/A+A/454/185 GROS M.: J/A+AS/109/447 Grosbol P.: J/A+A/423/849 Gross A.: J/other/APh/26.282 Grossenbacher R.: J/AJ/77/486 Grossi M.: J/MNRAS/342/738 J/MNRAS/346/787 J/MNRAS/355/1303 Grosso M.: III/204 J/A+A/408/305 J/A+AS/118/231 J/ApJ/573/359 Grosso N.: J/A+A/359/113 J/A+A/429/963 J/A+A/468/353 J/A+A/468/379 J/ApJS/160/319 J/ApJS/160/353

J/ApJS/160/511 Groth E.J.: J/ApJ/459/L27 J/ApJS/159/41 Grothues H.-G.: J/A+AS/121/247 J/A+AS/99/591 J/MNRAS/306/279 Grove R.: J/A+A/371/79 Groves B.A.: J/ApJ/619/755 Gruber D.: J/A+A/342/736 Gruber R.: IX/10A IX/11 IX/29 Gruberbauer M.: J/A+A/475/1053 Grueff G.: J/A+A/379/21 J/AJ/98/419 J/MNRAS/200/747 VIII/37 Gruendl R.A.: J/AJ/125/2239 J/AJ/127/477 Grullon S.: J/other/APh/26.282 Grundahl F.: J/A+A/465/965 J/AJ/127/1227 J/PASP/115/413 Grupe D.: J/A+A/350/805 J/AJ/127/156

Gruppioni C.: J/A+A/415/549 J/A+A/457/501 J/AJ/127/3075 J/MNRAS/302/222 J/MNRAS/305/297

J/MNRAS/325/1173 J/MNRAS/351/1290 J/MNRAS/358/397 J/MNRAS/371/1891 Gruppioni G: J/A+A/445/465 Grutzbauch R.: J/AJ/129/1832 J/MNRAS/364/146 Gry C.: J/A+A/303/137 GRY C.: J/A+AS/109/447 Gu M.: J/A+A/466/63 Gu M.F.: J/A+A/450/39 J/ApJS/156/105 Gu Q.: J/MNRAS/366/480 Gu S.: J/A+AS/125/475 Gualandi R.: J/A+A/458/225 Guanghua L.: J/A+AS/114/337 Guarcello M.G.: J/A+A/462/245 GUARINOS J.: II/135 Guarinos J.: II/149 GUARINOS J.: V/86 Guarnieri A.: J/PAZh/33/143 Guarnieri M. D.: J/A+AS/102/397 Guarnieri M.D.: J/A+A/331/70 J/A+AS/121/451 J/A+AS/137/75 Guarro J.: J/A+AS/125/399 Gubanov A.G.: J/PAZh/25/447 Gudas A.: J/BaltA/10/481 J/BaltA/11/153 Gudehus D.: J/ApJ/525/420 Gudmundsson E.H.: J/ApJ/416/276 Guedel M.: J/A+A/468/353 J/A+A/468/379 J/ApJ/484/394 Gueguen A.: J/MNRAS/382/279 Guelin M.: J/A+AS/142/181 Guenther E.: J/A+A/333/619 J/A+A/409/169 J/A+A/468/1039 J/AJ/121/1040 J/AJ/124/404 Guenther E.W.: J/A+A/466/1031 J/ApJ/632/L131 Guerrero A.: J/A+A/364/205 Guerrero M.A.: J/AJ/125/2239 J/AJ/127/477 Guertler J.: III/95 Guetter H.H.: J/AJ/105/1571 GUETTER H.H.: J/AJ/106/220 Guetter H.H.: J/AJ/127/3553 Guglielmetti F.: J/A+A/408/1029 GUGLIELMO F.: II/177 Guglielmo F.: J/A+A/334/609 J/A+AS/122/489 GUGLIELMO F.: J/A+AS/99/31 Gugliotti G. M.: V/114 Gugliotti G.M.: V/107 Guhathakurta P.: J/A+A/396/473 J/A+A/398/479 J/AJ/107/1745 GUHATHAKURTA P.: J/AJ/108/1786 Guhathakurta P.: J/AJ/112/2408 J/AJ/113/669 J/AJ/116/102 J/AJ/116/1757 J/AJ/124/234 J/AJ/124/3255

J/AJ/130/1693 J/AJ/131/332 J/ApJ/559/1060 J/ApJ/625/6 GUHATHAKURTA P.: J/ApJS/70/329 GUIBERT J.: J/A+A/299/696

Guibert J.: J/A+A/303/137 J/A+A/357/871 J/A+A/380/590 J/A+A/381/464 J/A+A/401/959

J/A+A/438/769 J/A+A/441/653 GUIBERT J.: J/A+AS/109/447

Guibert J.: J/A+AS/118/277 J/A+AS/120/275 J/A+AS/121/407 J/A+AS/123/507 J/A+AS/128/67

J/A+AS/131/239 J/A+AS/139/559 J/A+AS/145/11 Guichard J.: J/ApJS/142/79 Guiderdon B.: J/A+A/372/364 Guiderdoni B.: J/A+A/462/81 J/A+A/465/1099 GUIDERDONI B.: VI/47 VI/47 Guidorzi C.: J/A+A/474/793 J/ApJS/156/217 Guieu S.: J/A+A/446/485 Guijarro A.: J/A+A/399/457 Guilain C.: J/A+A/314/585

Guillout P.: J/A+A/318/111 J/A+A/454/301 J/A+A/476/1191 J/A+AS/122/201 J/A+AS/132/341

J/A+AS/146/323 Guimaraes M.M.: J/AJ/126/2971

Guinan E.F.: J/A+A/312/879 J/A+A/393/225 J/A+A/409/1017 J/A+A/410/671 J/A+A/463/1061

J/AJ/110/1350 J/AJ/128/846 J/ApJ/439/983 J/ApJ/484/394 Guinot B.: J/A+A/355/398 Gull T.: J/A+A/419/215 Gull T.R.: J/ApJ/612/481 J/ApJS/157/138 Gullieuszik M.: J/A+A/467/1025

Gulliver A.F.: J/A+A/306/924 J/A+A/367/597 J/A+A/394/187 J/A+A/400/1043 J/A+A/406/975

J/A+A/413/285 J/A+A/447/685 J/A+AS/129/563 Gullixson C.: II/95 Gulmen O.: J/A+A/409/959 GULYAEV A.P.: I/18 Gunasingha R.M.: J/other/APh/26.282 Gunn A.: J/PASJ/52/997

Gunn J.E.: III/88 J/A+A/371/833 J/A+A/423/755 J/AJ/107/1245 J/AJ/116/584 J/AJ/121/2308 J/AJ/121/2331 J/AJ/121/31 J/AJ/122/503 J/AJ/123/3409 J/AJ/123/567 J/AJ/125/1711 J/AJ/126/1023 J/AJ/126/1720 J/AJ/126/2579 J/AJ/127/3553 J/AJ/130/873 J/AJ/131/1 J/ApJ/241/486 J/ApJ/264/337 J/ApJ/306/30 J/ApJS/148/243 J/ApJS/155/257 J/ApJS/163/80 J/ApJS/78/1

Gunn K.: J/MNRAS/362/1371 J/MNRAS/369/156 Gunn K.F.: J/MNRAS/342/802 J/MNRAS/352/131 Gunnarsson C.: J/ApJ/561/106 Gupta A.C.: J/A+AS/142/373 J/A+AS/145/365 Gupta G.P.: J/ApJ/526/544 J/ApJS/123/295 Gupta R.: J/ApJS/152/201 J/ApJS/152/251 Gupta Ranjan: J/A+AS/141/175 Gupta V.: J/AJ/128/2965 Gursky H.: IX/3 Gurvits L.I.: J/A+A/342/378 J/AJ/130/2473 J/ApJ/616/110 J/ApJS/131/95 J/ApJS/155/33 J/PASJ/52/997 Gurwell M.: J/ApJS/73/661 Guseinov O.H.: J/PAZh/26/843 Gusenov O.H.: IX/8 Guseva N.G.: III/230 J/A+A/421/539 J/A+A/448/955 J/A+A/464/885 J/AJ/121/169 Gussie G.: J/AJ/114/2353

Gustafsson B.: J/A+A/275/101 J/A+A/327/587 J/A+A/350/955 J/A+A/353/287 J/A+A/425/187

J/A+A/440/321 J/A+AS/102/603 J/A+AS/129/237 J/ApJ/527/879 J/ApJ/530/783 Gustafsson M.: J/A+A/466/949 Guthrie B.N.G.: J/A+A/310/353 GUTHRIE B.N.G.: J/A+AS/93/255 Gutierrez C.M.: J/AJ/122/38 J/ApJ/602/664 J/ApJS/141/61 J/MNRAS/341/1066 Gutierrez P.J.: J/A+A/468/L13 Gutierrez-Soto J.: J/A+A/472/565 J/A+A/472/577 Gutzmer D.D.: J/A+A/371/79 Gutzwiller M.: II/230 Guyton K.M.: J/ApJ/525/420 Guziy S.: J/ApJ/641/L13 Guzman R.: J/A+A/464/L21 J/ApJ/481/673 J/ApJS/159/41 Guzzo L.: J/A+A/342/1 J/A+A/403/857 J/A+A/425/367 J/A+A/428/1043 J/A+A/457/79 J/A+AS/130/323

J/ApJ/514/148 J/MNRAS/274/1071 J/MNRAS/276/689 GUZZO L.: VII/175 Gwinn C.R.: J/ApJ/393/149 J/ApJ/429/253 Gwyn S.: J/A+A/403/857 J/MNRAS/371/221 Gwyn S.D.J.: J/ApJ/633/630 G.Zhao.: J/A+A/371/943

H.: J/A+A/448/983 Haarsma D. B.: J/ApJ/510/64 Haarsma D.B.: J/AJ/130/1977 Haas M.: J/A+A/401/873 J/A+A/456/953 J/A+A/459/55 J/A+A/473/369 Haas M.R.: J/ApJS/162/346 Haase J.: J/A+A/434/53 J/A+A/454/423 J/A+A/478/83 Haase S.: J/ApJS/115/1

Haberl F.: IX/10A IX/11 IX/29 J/A+A/318/111 J/A+A/344/521 J/A+A/373/438 J/A+A/426/11 J/A+A/434/483 J/A+A/448/1247 J/A+A/451/835 J/A+A/465/375 J/A+AS/122/201 J/A+AS/132/341 J/A+AS/136/81 J/A+AS/139/277 J/A+AS/142/41 J/A+AS/143/391 J/A+AS/147/75

Haberzettl L.: J/A+A/441/905 J/A+A/458/341 Habgood M.J.: J/A+A/415/521 Habing H.: J/A+A/267/515 J/A+A/349/L69 J/ApJ/552/289 Habing H.J.: II/228A J/A+A/348/897 J/A+A/362/215 J/A+A/365/545 J/A+A/366/481 J/A+A/377/945

J/A+A/393/115 J/A+A/406/51 J/A+A/418/103 J/A+A/429/837 J/A+A/435/575 J/A+A/442/165 J/A+AS/122/79 J/A+AS/124/509 J/A+AS/128/35 J/A+AS/90/327 J/A+AS/92/43

HABING H.J.: J/A+AS/93/121 Habing H.J.: J/MNRAS/376/313 HABING H.J.: VII/109 Habing H.J.: VIII/7A Habing R.J.: J/A+AS/119/333 Hachisuka K.: J/PASJ/52/631 Hack M.: J/A+AS/124/425 Hacking P.: II/156A Hackman T.: J/A+A/350/513 J/A+A/362/223 J/A+A/383/197 J/A+A/476/881 J/A+AS/139/513 Hackwell J.A.: II/95 Hadfield L.J.: J/A+A/439/265 Hadjifotinou K.G.: J/A+A/391/775

Hadrava P.: J/A+A/309/521 J/A+A/312/879 J/A+A/319/867 J/A+A/324/965 J/A+A/328/551

J/A+A/408/611 J/A+A/419/607 J/A+A/446/583 J/A+AS/125/75 Haehnelt M.G.: J/AJ/124/2955 HAENSEL P.: J/A+A/291/155 J/A+AS/108/455 Haffner L.M.: II/249 Hagan L.: VI/74 Hagele G.F.: J/MNRAS/355/1237 Hagen H.-J.: J/A+A/366/771 J/A+A/374/907 Hagen H.J.: J/A+A/378/556 Hagen H.-J.: J/A+A/397/463 J/A+A/400/939 J/A+A/406/535 J/A+A/442/109 J/A+A/449/425 HAGEN H.-J.: J/A+AS/110/469 Hagen H.J.: J/A+AS/116/43

Hagen H.-J.: J/A+AS/128/507 J/A+AS/134/483 J/A+AS/135/511 J/A+AS/137/299 J/A+AS/142/417

VI/116 Hagen-Thorn A.V.: J/PAZh/30/243 Hagen-Thorn E.I.: III/201 J/PAZh/32/672 Hagen-Thorn V.A.: J/A+A/385/55 J/PAZh/30/243 Haggkvist L.: II/21A Haig D.: J/ApJ/638/293 Haikala L.: J/A+A/372/276 Haikala L.K.: J/A+A/421/1087 J/A+AS/127/1 Haiman Z.: J/A+A/278/36 J/AJ/131/49 Hainaut O.: J/A+AS/125/399 Hainaut O.R.: J/A+A/380/347 J/A+A/389/641 J/A+A/406/593 J/A+AS/112/125 Hainaut-Rouelle M.-C.: J/A+AS/112/125 Haines C.: J/A+A/424/79 Haines C.P.: J/MNRAS/358/1116 Haisch B.M.: J/ApJ/348/253 Haisch K.E.Jr: J/AJ/131/527 Haissinski J.: J/A+A/389/149 J/A+A/454/185 Hakala P.: J/MNRAS/373/687 Hakala P.J.: J/A+AS/134/287 Hakala T.: J/AJ/133/1947 Hakkila J.: IX/20A Hakopian S.A.: J/A+A/414/487 Halbwachs J.L.: I/121 I/207 Halbwachs J.-L.: J/A+A/282/831 J/A+A/373/576 J/A+AS/137/451 Halbwachs J.-L. Mayor M.: J/A+A/314/846 Halbwachs J.L.: V/53A Hale A.: J/PASJ/52/997 Hale J.A.: IX/13 Hales S.E.G.: J/MNRAS/376/1251 HALES S.E.G.: VIII/21 VIII/22 VIII/23 VIII/24 VIII/25 Hales S.E.G.: VIII/31 Halkola A.: J/MNRAS/372/1425 Hall D.M.: I/232 I/268 I/289 J/AJ/110/880 J/AJ/120/1120 J/AJ/122/1586 J/AJ/124/2254 J/AJ/127/539 Hall D.N.B.: III/114 Hall D.S.: J/A+A/312/879 J/A+A/362/223 J/A+A/383/197 HALL D.S.: J/AJ/105/2265 J/AJ/107/1868 V/71 Hall D.S.: V/76 Hall J.C.: J/PASP/108/313

Hall P.B.: J/AJ/114/2269 J/AJ/121/1840 J/AJ/121/2843 J/AJ/123/567 J/AJ/124/1651 J/AJ/125/1711 J/AJ/126/1023 J/AJ/126/1720 J/AJ/126/2125 J/AJ/126/2209 J/AJ/126/2579 J/AJ/129/2542 J/AJ/130/1345 J/AJ/130/873 J/AJ/131/1 J/AJ/131/84 J/AJ/134/1103 J/ApJ/462/614 J/ApJ/565/208 J/ApJS/104/185 J/ApJS/119/189 J/ApJS/129/475 J/ApJS/165/1 VII/243

Hall P.J.: J/A+AS/127/185 Hall R.T.: II/53 Hallet N.: J/A+A/430/373 J/A+A/465/71 J/A+AS/130/333 VII/238 Hallgren A.: J/other/APh/26.282

Halliday C.: J/A+A/427/397 J/A+A/444/365 J/ApJ/630/206 J/MNRAS/326/473 J/MNRAS/369/120

J/MNRAS/369/97 J/MNRAS/371/1777 Halpern J.P.: J/ApJ/599/886 J/ApJS/150/181 Halpern M.: J/ApJS/148/97 J/MNRAS/372/1621 J/MNRAS/376/1073 Halzen F.: J/other/APh/26.282 Hamabe M.: J/ApJS/85/249 J/MNRAS/357/1357 J/PASJ/55/1079 VII/147 Hamadache C.: J/A+A/454/185 Hamaguchi K.: J/ApJ/567/423 J/ApJ/607/904 J/ApJ/635/214 Hamami I.: J/MNRAS/316/514 Hamann W.-R.: J/A+A/437/467 Hambaryan V.: J/A+A/353/1044 Hambaryan V.V.: J/other/Ap/41.356

Hambly N.C.: J/A+A/347/69 J/A+A/373/608 J/A+A/375/863 J/A+A/397/575 J/A+A/416/125

J/A+A/421/763 J/A+A/435/363 J/A+A/468/163 HAMBLY N.C.: J/A+AS/100/607 J/A+AS/109/29

Hambly N.C.: J/A+AS/139/231 J/AJ/129/413 J/AJ/130/1658 J/ApJS/134/103 J/MNRAS/313/347

J/MNRAS/362/689 J/MNRAS/373/95 J/MNRAS/374/372 J/MNRAS/380/712 J/other/Sci/292.698 Hamdani S.: J/A+A/360/509 Hamdy M.A.: J/other/Ap+SS/203.53 Hameed S.: J/MNRAS/376/960 Hameury J.M.: J/A+AS/123/273 Hamilton D.: J/A+A/320/757 Hamilton J-C.: J/A+A/389/149 Hamlin D.W.: II/68 Hammarback G.: J/A+A/151/372

Hammer F.: J/A+A/423/867 J/A+A/430/115 J/A+A/435/507 J/A+A/455/107 J/A+A/462/535

J/A+A/462/81 J/A+A/465/1099 J/ApJ/481/49 J/ApJ/499/112 J/ApJ/517/148 VII/225A Hammer H.: J/ApJ/464/79 Hammersley P.: J/ApJ/491/L31 Hammersley P.L.: II/270 Hammond J.C.: I/100A Hamuy M.: II/179 J/AJ/112/2408 Hamzaoglu E.: J/MNRAS/355/307 Han F.: J/A+AS/127/181 J/other/PPMtO/14.185 Han I.: J/AJ/131/1015 J/ApJ/634/451 J/PASP/118/636 Han J.L.: J/A+AS/136/571 Han J.-L.: J/other/AcApS/20.366 J/other/ChJAA/6.237 Han K.: J/other/APh/26.282 Han M.: J/AJ/113/1001 J/ApJ/507/655 J/ApJ/508/491 J/ApJ/512/48 J/ApJ/521/155 J/ApJ/523/540 Han W.: J/AJ/126/2840 Han Z.: J/A+A/430/223 J/A+A/455/247 J/A+A/471/795 J/MNRAS/319/215 J/MNRAS/380/1098 Hanada H.: J/PASJ/52/631 Hancock M.: J/AJ/125/1696 J/AJ/127/1360 Hancock P.J.: J/MNRAS/371/898 Handa T.: J/ApJS/100/125 Handler G.: J/A+A/374/1056 J/A+AS/135/493 J/ApJS/158/193 Hands A.D.P.: J/MNRAS/351/31 Hanes D.A.: J/AJ/123/2490 J/ApJ/559/812 J/MNRAS/355/608 Hanisch R.J.: J/AJ/130/2439 J/ApJ/612/437 J/ApJ/616/1284 HANISCH R.J.: VIII/29 Hanish D.J.: J/ApJS/165/307 Hanley C.J.: J/A+A/330/515 Hanlon L.: J/A+A/385/377 Hanner M.S.: J/A+AS/127/1 HANNI M.: VI/60 Hannikainen D.: J/A+A/467/585 Hanninen J.: J/A+A/423/517 Hansen B.M.S.: J/AJ/127/2771 Hansen K.H.: J/ApJ/586/1356 HANSEN L.: J/A+A/279/61 Hansen L.: J/A+A/388/1 J/A+A/409/439 J/A+A/435/781 HANSEN L.: J/A+AS/104/179 HANSEN O.L.: J/AJ/80/1011 Hansen S.: J/ApJS/148/243 Hanski M.: J/A+A/427/769 J/A+A/430/373 J/A+A/465/71 J/A+AS/132/305 J/A+AS/143/357 Hanski M.O.: J/A+A/372/276 Hanslmeier A.: J/A+A/447/735 Hanslmeier: J/A+A/390/707 HANSON G.J.: J/AJ/107/1453 Hanson K.: J/other/APh/26.282 Hanson M.M.: J/ApJS/107/281 J/ApJS/161/154 J/MNRAS/308/473 Hanson R.B.: I/199A I/200 I/214 I/283A I/293 J/AJ/119/2866 HANSON R.B.: J/AJ/80/379 Hanuschik R.W.: J/A+A/407/1157 Hao C.N.: J/ApJ/625/78 Hao L.: J/AJ/126/1720 J/AJ/126/2125 Hara H.: I/249 Hara T.: J/PASJ/52/631 Harada N.: J/PASJ/52/631 HARAMUNDANIS K.L.: I/111 II/6 Harbeck D.: J/A+A/433/751 J/AJ/125/197 J/AJ/127/1588 J/AJ/127/704 J/ApJ/633/844 Harbeck D.R.: J/AJ/131/895 Hardcastle M.J.: J/A+A/427/387 Harden F.R. Jr: J/ApJ/414/846 Harder S.: J/A+AS/129/337 Hardin D.: J/A+A/389/149 Harding P.: J/AJ/120/2496 J/AJ/125/2502 J/ApJ/508/491 J/ApJ/512/48 J/ApJ/521/155 J/ApJ/523/540

J/ApJ/616/804 Hardorp J.: III/15 III/15 III/15 III/76A III/76A III/76A Hardtke D.: J/other/APh/26.282 Hardtke R.: J/other/APh/26.282 Hardy E.: J/A+AS/61/473 J/AJ/118/1468 J/AJ/118/1489 J/AJ/127/840 J/MNRAS/279/595 Harenberg T.: J/other/APh/26.282 Hargis J.R.: J/ApJ/608/243 J/ApJ/628/729 Harju J.: J/A+AS/132/211 Harjunpaa P.: J/A+A/421/1087 Harker J.: J/ApJS/157/175 HARMANEC P.: J/A+A/294/135

Harmanec P.: J/A+A/304/235 J/A+A/309/521 J/A+A/312/879 J/A+A/319/867 J/A+A/324/965 J/A+A/328/551 J/A+A/337/183 J/A+A/367/236 J/A+A/387/580 J/A+A/405/1087 J/A+A/408/611 J/A+A/415/289 J/A+A/416/669 J/A+A/419/607 J/A+A/440/249 J/A+A/446/583 J/A+A/455/1037 J/A+A/455/259 J/A+A/459/849 J/A+A/463/1061 J/A+A/463/233 J/A+A/464/263

HARMANEC P.: J/A+AS/107/403 Harmanec P.: J/A+AS/125/75 Harmer D.: J/A+A/364/102 Harmon B.A.: J/ApJS/154/585 Harnden F.R.: J/ApJ/348/557 J/ApJ/351/492 J/ApJS/102/75 Harnden F.R.Jr: J/A+A/355/651 Harnden F.R.JR: J/A+A/421/175 Harnden F.R.Jr.: J/A+A/460/133 Harnden F.R.Jr: J/ApJ/315/687 J/ApJ/325/798 J/ApJ/348/253 J/ApJ/582/382 J/ApJ/588/1009 Harnden F.R.JR: J/ApJ/606/466 Harnden F.R.Jr: J/ApJ/608/781 J/ApJS/150/19 Harnden F.R.JR: J/ApJS/160/319 J/ApJS/160/450 Harnden: J/A+A/341/751 Harnden Jr. F.R.: IX/13 Haro G.: II/131 III/74 Harold G. Corwin Jr: VII/239A Harper D.: J/A+A/391/775 J/A+A/422/377 J/A+AS/121/65 J/A+AS/136/257 Harries T.J.: J/A+A/468/1009 J/MNRAS/353/601 HARRINGTON R.: I/169 Harrington R.: I/270 Harrington R.S.: I/157 J/AJ/105/1571 Harris A.W.: J/A+A/365/110 J/A+A/378/116 J/A+A/379/564 VII/108C Harris D.E.: IX/13 Harris E.L.: J/A+A/364/712 Harris G.J.: J/MNRAS/357/471 VI/119 VI/121 HARRIS G.L.H.: J/AJ/106/1533 Harris G.L.H.: J/AJ/115/1801 J/AJ/128/712 J/AJ/128/723 J/ApJ/559/812 Harris H.C.: J/AJ/103/638 J/AJ/105/1196 J/AJ/105/1571 HARRIS H.C.: J/AJ/106/604 Harris H.C.: J/AJ/123/3409 J/AJ/124/1651 J/AJ/126/1023 J/AJ/127/2838 J/AJ/127/3034 J/AJ/128/426

J/AJ/129/2542 J/AJ/131/1674 J/AJ/131/582 J/ApJ/607/426 J/ApJS/155/257 HARRIS H.C.: V/48

Harris J.: J/AJ/121/891 J/AJ/122/3046 J/AJ/123/855 J/AJ/127/1531 J/AJ/128/1606 J/AJ/131/912

J/ApJ/603/503 Harris M.J.: J/ApJS/156/69 Harris P.: J/other/APh/26.367

Harris W.E.: J/AJ/115/1801 J/AJ/117/247 J/AJ/128/712 J/AJ/128/723 J/ApJ/383/487 J/ApJ/462/1

J/ApJ/599/1067 HARRIS W.E.: J/ApJS/73/671 Harris W.E.: J/PASP/99/739 VII/195 VII/202 Harrison F.A.: J/ApJ/596/944 J/ApJS/156/35 J/ApJS/165/19 Harrison P.A.: J/A+A/472/805 Harrison T.E.: J/ApJ/595/187 J/ApJ/602/264 Hart J.E.: J/other/APh/26.282 Hartig G.: J/ApJ/621/53 Hartig G.F.: J/AJ/128/1990 J/ApJ/612/202 J/ApJS/150/1 Hartigan P.: J/AJ/112/2408 J/AJ/124/445 J/AJ/127/3456 Hartkopf: V/111

Hartkopf W.I.: B/sb9 J/AJ/106/1639 J/AJ/106/352 J/AJ/106/637 J/AJ/108/2299 J/AJ/111/370 J/AJ/111/393 J/AJ/111/936 J/AJ/117/1890 J/AJ/119/3084 J/AJ/120/1120 J/AJ/121/3224 J/AJ/122/1586 J/AJ/123/1023 J/AJ/124/2254 J/AJ/127/539 J/AJ/128/3012 J/ApJ/525/420 V/122 V/33 HARTKOPF W.I.: VI/24 HartkopF W.I.: VI/46 HARTL H.: J/A+AS/93/553 Hartlap J.: J/A+A/470/821 Hartley M.: J/MNRAS/362/689 J/other/PASA/17.92 V/127 VII/191 Hartman H.: J/A+A/419/215 Hartman J. D.: J/A+A/466/743 Hartman J.D.: J/AJ/128/1761 J/AJ/129/1596 J/ApJ/626/L5 J/MNRAS/371/1405

Hartman R.C.: IX/9 J/ApJ/481/95 J/ApJS/101/259 J/ApJS/107/227 J/ApJS/123/79 J/ApJS/135/155

J/ApJS/94/551 Hartmann D.: VIII/54 VIII/76 Hartmann D.H.: J/ApJ/446/115 Hartmann L.: J/A+A/440/595 J/AJ/129/856 J/AJ/129/907 J/AJ/131/1574 J/AJ/131/527 J/ApJ/578/914

J/ApJ/638/897 J/ApJS/101/117 Hartmann L.W.: J/AJ/128/805 J/AJ/129/363 J/AJ/130/188 J/ApJS/91/625 Hartner G.: IX/10A IX/29 HARTWICK F.D.A.: J/AJ/104/1706 Hartwick F.D.A.: J/AJ/121/1840 J/AJ/96/1632 HARTWICK F.D.A: VII/143 Harutyunyan A.: J/MNRAS/369/1880 Harvanek M.: J/A+A/444/643 J/AJ/123/3409 J/AJ/131/1674 J/AJ/132/676 J/PASJ/55/771 Harvey P.: J/ApJ/638/293 Harwit M.: J/A+A/372/364 HARWOOD D.N.: I/168 Haschick A.D.: J/ApJ/463/205 Hasegawa T. and Nishihara E.: J/A+A/365/409 Hasegawa T.: J/ApJS/100/125 J/MNRAS/316/326 J/PASJ/56/295 Hasegawa W.: J/PASJ/52/631 Haser S.M.: J/A+A/321/531 HASHIMOTO O.: J/A+AS/107/445 Hashimoto Y.: J/A+A/476/1191 J/ApJS/130/237 Hashmall J.: V/77

Hasinger G.: II/208 II/253 IX/10A IX/29 IX/32 J/A+A/278/36 J/A+A/309/116 J/A+A/317/328

J/A+A/329/482 J/A+A/371/833 J/A+A/373/63 J/A+A/393/425 J/A+A/398/901 J/A+A/444/79 J/A+A/451/457 Hasinger G: J/A+A/466/41

Hasinger G.: J/A+A/476/1191 J/A+AS/114/215 J/ApJ/595/685 J/ApJS/139/369 J/ApJS/155/271

J/ApJS/155/73 J/ApJS/161/21 J/MNRAS/311/456 J/PASJ/53/445 Haslam C.G.T. et al: VIII/43 Haslam C.G.T.: J/A+AS/58/197 VIII/8A Hassall B.J.M.: J/A+A/476/1191 J/MNRAS/274/1194 Hassan: J/A+A/372/879 Hassan S.M.: J/A+AS/115/325 Haswell C.A.: J/A+A/467/785 J/MNRAS/372/1117 J/MNRAS/379/647 J/MNRAS/379/816 Hatchell J.: J/A+A/440/151 J/A+A/453/1003 J/A+A/468/1009 J/A+A/472/187 J/A+AS/133/29 Hathi N.P.: J/ApJ/639/724 Hatzes A.P.: J/AJ/127/3579 J/ApJ/632/L131 Hatzidimitriou D.: J/A+A/426/11 J/A+A/448/1247 J/A+A/451/835 J/A+AS/113/539 J/A+AS/122/507

J/MNRAS/289/225 J/MNRAS/341/534 J/MNRAS/355/1196 VII/183 Hatziminaoglou E.: J/A+A/379/436 J/A+A/384/81 J/A+A/392/741 J/A+A/434/475 J/A+A/458/245

J/A+A/459/693 J/AJ/129/1198 J/AJ/129/2074 J/MNRAS/351/1290 Hau G.K.T.: J/MNRAS/277/125 HAUCK B.: II/107 II/128 II/158A II/167 Hauck B.: II/215 II/57 II/74 HAUCK B.: III/107 III/163 III/165 VI/5 HAUG U.: J/A+AS/106/79 HAUK B.: III/56 III/96 Hauschildt M.: VIII/68 Hauschildt P.: J/A+A/337/403 Hauschildt P.H.: J/A+A/439/1137 J/A+A/442/281 J/MNRAS/357/945 VI/125 Hauschildt T.: J/other/APh/26.282 Hauser M.G.: J/A+AS/127/1 Havlen R.: J/AJ/112/2408 Havrilova N.V.: J/AZh/82/437 Hawarden T.G.: J/MNRAS/373/781 Hawkins E.: J/MNRAS/350/1485 Hawkins I.: J/AJ/107/751 J/ApJS/94/127 HAWKINS M.R.S.: I/185 J/A+A/296/665 Hawkins M.R.S.: J/A+A/386/97 J/A+A/424/519 HAWKINS M.R.S.: J/A+AS/100/607 J/A+AS/109/171 J/A+AS/109/29 Hawkins M.R.S.: J/A+AS/140/351 J/A+AS/143/465 HAWKINS M.R.S.: J/MNRAS/260/202 Hawkins M.R.S.: J/MNRAS/274/1194 J/MNRAS/275/1102 VII/221 Hawley S.: J/AJ/126/1023 Hawley S.L.: III/198 J/A+A/330/515 J/A+AS/126/81 J/A+AS/143/409 J/AJ/123/3356 J/AJ/123/3409

J/AJ/124/1118 J/AJ/126/3007 J/AJ/128/426 J/AJ/128/463 J/AJ/129/2428 J/AJ/130/1871 J/AJ/131/1674 J/AJ/133/439 J/ApJ/607/426

Hawley W.N.: J/ApJS/99/551 Hayashi H.: J/ApJS/102/411 Hayashi M.: J/AJ/131/1608 J/ApJS/100/125 Hayashi S.S.: J/AJ/131/1608 Hayashida K.: IX/19 Hayashino T.: J/PASJ/57/165 J/PASJ/58/313 Hayes J.J.E.: J/AJ/124/1651 Haynes M.P.: J/AJ/105/1251 J/AJ/105/1271 J/AJ/113/1197 J/AJ/113/22 J/AJ/116/2728 J/AJ/116/2738

J/AJ/116/2805 J/AJ/117/1668 J/AJ/117/2039 J/AJ/118/1468 J/AJ/118/1489 J/AJ/127/3273 J/AJ/129/2119 J/AJ/130/1037 J/AJ/130/2613 J/AJ/131/716 J/ApJ/636/214 VIII/77

Haynes R.F.: J/A+A/171/261 J/A+AS/111/311 J/A+AS/120/77 J/A+AS/121/321 J/A+AS/127/119 J/A+AS/130/421 J/A+AS/90/327 J/AJ/124/1954 J/AJ/128/16 J/MNRAS/346/787 J/MNRAS/355/1303 VIII/20 VIII/56

Hays D.: J/other/APh/26.282 Hayward T.L.: J/ApJ/635/280 Haywood M.: J/MNRAS/325/1365 Hazard C.: J/AJ/87/739 J/AJ/90/1437 VIII/61 Hazen M.: J/A+A/323/113 HAZEN M.L.: J/AJ/107/1764 Hazen M.L.: J/other/IBVS/5298 J/PAZh/28/201 He J.: J/A+AS/115/283 He J.H.: J/A+A/434/201 J/other/NewA/9.509 He J.-Z.: J/A+A/474/783 He L.: J/other/SAAOC/15.73 He X.-T.: J/AJ/123/578 J/other/AcApS/20.366 He Y.P.: J/A+AS/100/243 Heafield K.: J/AJ/130/825 Healey S.E.: J/ApJ/626/95 Heap S.: J/ApJ/612/481 Heap S.R.: J/AJ/107/1054 J/AJ/116/941 J/ApJ/636/631 J/ApJS/146/209 Hearnshaw J.B.: J/MNRAS/330/137 J/MNRAS/331/609 Hearty T.: J/A+A/341/163 J/A+A/353/1044 J/AJ/127/2659 J/AJ/128/787 Heasley J.N.: J/MNRAS/238/1171 Heathcote S.R.: II/179 Hebb L.: J/MNRAS/370/954 J/MNRAS/377/741

Heber U.: III/137 J/A+A/378/556 J/A+A/400/877 J/A+A/400/939 J/A+A/430/223 J/A+A/442/1023

J/A+A/447/173 HEBER U.: J/A+AS/103/445 Heck A.: III/83 VI/37 HECK A.: VI/45 VI/8 Heckathorn H.M.: II/92 Heckert P.A.: J/A+A/420/595 J/A+A/424/993 J/A+AS/139/513 Heckman T.: J/A+A/431/793 J/AJ/125/1817 Heckman T.M.: J/AJ/123/567 J/AJ/126/2125 J/AJ/126/2579 J/AJ/127/213 J/AJ/128/1002 J/AJ/130/825

J/ApJ/602/231 J/ApJ/619/L95 J/ApJS/105/93 J/ApJS/144/21 J/ApJS/164/38 J/ApJS/165/307 J/ApJS/99/173 J/MNRAS/362/9

Heckmann O.: I/61B I/69 J/AN/283/109 Hecquet J.: J/A+A/421/509 Heemskerk M.H.M.: J/A+A/473/523 Heeschen D.S.: J/ApJS/13/65 Heeter R.F.: J/A+A/416/383 J/A+A/424/363 J/A+A/433/745 J/A+A/455/1157 Heflin M.: J/ApJS/75/801 VIII/52B Heflin M.B.: VIII/52B Heidt J.: J/A+A/368/16 J/A+A/398/49 J/A+A/418/885 J/A+A/420/97 J/A+A/421/41 J/A+A/422/907 Heikkila J.: J/A+A/427/769 Heiles C.: II/226 J/AJ/87/1165 J/ApJS/145/329 J/ApJS/151/271 VIII/7A VIII/7A VIII/8A VIII/8A

VIII/9 Heim T.: J/A+A/323/337 Heinaemaeki P.: J/A+A/459/L1

Heinamaki P.: J/A+A/462/811 J/A+AS/128/459 J/A+AS/129/577 J/A+AS/130/109 J/A+AS/132/305

J/A+AS/143/357 J/AN/327/365 Heinamaki P. Valtonen M.: J/A+AS/124/559 Heindl W.A.: J/A+A/395/129 J/A+A/407/1039 Heinize K.G.: III/86 Heinke C.: J/other/Sci/292.2290 Heinke C.O.: J/ApJ/625/796 Heinrichsen I.: J/AJ/131/250 Heinsheimer T.F.: II/96 HEINTZ W.D.: I/19 Heintz W.D.: J/AJ/108/2338 J/ApJS/105/475 J/ApJS/117/587 J/ApJS/58/439 J/ApJS/74/275 HEINTZ W.D.: J/ApJS/83/351 Heintz W.D.: J/ApJS/99/693 V/39 Heinzelmann G.: J/A+A/390/39 J/A+A/406/L9 J/A+A/410/813 Heise J.: J/other/APh/26.282 Heisler C.A.: J/ApJS/132/37 Heiter U.: J/A+A/373/625 J/A+A/405/1095 J/A+AS/123/93 J/AJ/129/1063 Heithausen A.: J/A+AS/144/123 VIII/66 VIII/66 Hekker S.: J/A+A/475/1003 Helbig P.: J/A+A/318/680 J/MNRAS/377/430 Helbing K.: J/other/APh/26.282 Held E.V.: J/A+A/315/40 J/A+A/431/1189 J/A+A/452/179 J/A+A/467/1025 J/MNRAS/363/734

Helfand D.J.: IX/18 J/AJ/112/407 J/AJ/121/1872 J/AJ/127/2565 J/AJ/129/348 J/AJ/130/156 J/AJ/130/586 J/ApJ/461/127 J/ApJ/596/944 J/ApJ/607/60 J/ApJS/126/133 J/ApJS/135/227 J/ApJS/143/1 J/ApJS/156/35 J/ApJS/165/19 J/ApJS/74/181 J/ApJS/91/347 J/MNRAS/351/31 VIII/48 VIII/51 VIII/59 VIII/71

Heller: J/A+AS/135/145 Hellier C.: J/A+A/467/785 J/MNRAS/372/1117 J/MNRAS/379/647 J/MNRAS/379/816 Helling C.: J/A+A/361/641 Hellwig M.: J/other/APh/26.282 Helmboldt J.F.: J/ApJ/613/914 Helmer G.: J/A+AS/110/107 J/A+AS/125/399 Helmer L.: I/282 I/99 I/99 I/99 I/99 J/A+A/312/1031 Helmi A.: J/A+A/459/423 J/AJ/125/2502 Helmich F.P.: J/A+AS/124/205 Helou G.: J/A+A/424/371 J/AJ/125/2411 J/AJ/131/250 J/ApJ/633/871 J/ApJS/73/359 J/PASP/101/360

VII/73 Helou: J/PASP/110/779 Helt B.E.: J/A+A/392/215 J/A+A/461/1065 HELT B.E.: J/A+AS/101/563 J/A+AS/112/123 Helt B.E.: J/A+AS/115/315 J/A+AS/125/471 HEMENWAY P.: I/169 Hempel H.: J/A+A/435/65 Hempel M.: J/A+A/415/123 Henden A.: J/A+A/372/145 J/A+A/434/1107 J/A+A/445/617 J/A+AS/143/343 J/other/IBVS/5526 Henden A. Munari U.: J/A+A/458/339

Henden A.A.: II/230 J/AJ/111/902 J/AJ/112/2757 J/AJ/115/296 J/AJ/127/3553 J/PASP/107/324

J/PASP/109/441 Henderson C.P.: J/ApJ/635/280 Henderson D.: J/AJ/131/250 Hendrickson M.: J/AJ/127/3553 Hendry M.A.: J/A+A/411/361 J/A+A/437/467 Henelius N.: J/A+AS/132/305 Henize K.G.: J/ApJS/2/315 Henkel B.: J/ApJ/635/950 Henkel C.: J/A+A/305/960 J/A+A/361/1079 J/A+A/394/975 J/A+A/430/549 Hennawi J.: J/AJ/127/3553 Hennawi J.F.: J/AJ/131/1 Hennemann M.: J/A+A/474/883 Hennessy G.: J/AJ/124/1651 Hennessy G.S.: J/AJ/120/1120 J/AJ/121/31 J/AJ/122/1586 J/AJ/122/503 J/AJ/123/1807 J/AJ/123/3409

J/AJ/123/567 J/AJ/124/2254 J/AJ/125/1817 J/AJ/126/1023 J/AJ/126/2579 J/AJ/127/3034 J/AJ/127/539 Hennig L.A.: J/ApJ/612/511 Henning P.A.: J/A+A/391/887 J/AJ/123/873 J/AJ/124/1954 J/AJ/124/690 J/AJ/128/16 J/AJ/129/220

J/AJ/132/242 J/MNRAS/337/641 Henning T.: J/A+A/291/943 J/A+A/318/879 J/A+A/423/983 J/A+A/434/987 J/A+A/449/583 J/A+A/474/883 J/A+AS/115/285 J/A+AS/124/385 J/AJ/129/1049 J/ApJS/155/123 J/ApJS/161/361 J/ApJS/166/549 J/PASP/118/1690

Henrichs H.F.: J/A+A/383/580 J/A+A/409/275 J/A+A/411/565 J/A+AS/105/305 J/A+AS/116/257 Henriksen M.: J/ApJ/511/595 Henry G.W.: J/A+A/362/223 J/A+A/383/197 J/A+A/402/1033 HENRY G.W.: J/AJ/105/2265 Henry G.W.: J/AJ/129/1669 J/AJ/129/2026 J/AJ/129/2815 J/AJ/130/794 J/ApJ/634/625 Henry J.P.: J/A+A/278/379 J/A+A/419/47 J/A+A/461/977 HENRY J.P.: J/ApJ/365/66 Henry J.P.: J/ApJ/594/154 J/ApJ/602/571 J/ApJ/617/192 J/ApJ/622/187 J/ApJS/149/29 J/ApJS/162/304 Henry R.C.: II/269 II/42 II/43 II/44 Henry S.M.: J/AJ/129/2815 Henry T.: J/A+A/325/159 Henry T.J.: J/AJ/106/773 J/AJ/111/439 J/AJ/129/413 J/AJ/130/1658 Hensberge H.: J/A+A/358/553 J/A+AS/128/409 J/MNRAS/356/545 Hensler G.: J/A+A/423/517 Henstock D.R.: J/A+A/472/763 J/ApJS/100/1 J/ApJS/95/345 Henze W.: IX/20A J/ApJS/154/585 Herald D.: I/291 Heras A.: J/A+A/357/839 J/A+A/390/1033 Heraudeau: J/A+AS/136/509 Heraudeau P.: J/A+AS/117/417 J/A+AS/118/111 J/A+AS/133/317 J/AJ/127/3075 J/MNRAS/351/1290

VII/206 Herbert-fort S.: J/ApJ/635/123 Herbert-Fort S.: J/PASP/116/622 J/PASP/118/1077 Herbig G.H.: J/AJ/123/304 J/AJ/128/1233 J/AJ/131/1530 J/ApJ/497/736 V/73A Herbig T.: J/ApJS/81/83 Herbst E.: J/A+AS/119/111 J/ApJ/499/517 J/ApJ/500/1059 J/ApJ/519/697 J/ApJS/162/428 Herbst T.M.: J/AJ/129/1534 Herbst W.: J/A+A/396/513 J/A+A/417/557 J/A+A/430/1005 J/A+A/461/183 HERBST W.: J/AJ/107/1453 Herbst W.: J/AJ/118/1043 J/AJ/127/1602 J/ApJS/160/319 J/ApJS/160/450 VII/218 Herbstmeier U.: J/A+A/401/873 J/MNRAS/351/1290 Herczeg G.J.: J/ApJS/165/256 Hergenrother C.W.: J/AJ/131/527 Hering R.: J/A+AS/122/529 V/67 Hermann G.: J/A+A/390/39 J/A+A/406/L9 J/A+A/410/813 Hermsen W.: J/A+AS/143/145 Hernan-Caballero A.: J/AJ/129/1198 Hernandez E.B.: III/58 Hernandez J.: J/AJ/127/1158 J/AJ/129/363 J/AJ/129/856 J/AJ/129/907 J/AJ/130/188 Hernandez M.: J/AJ/112/2408 Hernandez M.M.: J/A+A/346/586 Hernandez O.: J/A+A/466/905 Hernandez-Toledo H.: J/A+A/412/633 Hernandez-Toledo H.M.: J/A+A/379/54 J/AJ/129/682 Hernius O.: J/A+AS/115/295 Herpin F.: J/MNRAS/367/1609 Herquet P.: J/other/APh/26.282 Herranz D.: J/ApJ/621/53 Herrera D.: J/AJ/127/1158 J/AJ/134/1103 J/ApJ/621/104 J/ApJS/162/1 Herrera: J/A+A/356/529 Herrero A.: J/A+A/437/467 J/A+A/457/265 Herrero J.L.: J/AJ/121/66 J/AJ/123/1292 Herrmann K.A.: J/ApJ/614/167 Herrnstein R.M.: J/AJ/127/3399 Hershey J.L.: J/PASP/110/1012 Hertenstein T.: J/A+AS/114/175 Herter T.: J/AJ/113/1197 J/AJ/113/22 J/AJ/117/1668 J/AJ/127/3273 J/ApJS/146/267 HERTLING G.: J/A+AS/104/1 Hertling G.: J/AJ/107/2093 J/ApJS/126/1 VII/208 Hertz P.: J/A+AS/134/221 HERTZ P.: J/ApJS/93/145 Herwig F.: J/A+A/339/123 J/ApJS/155/651 HERZ N.: I/12 Herzog E.: VII/4A Hes R.: J/A+A/319/757 HESKE A.: J/A+AS/93/121 Hess T.R.: J/AJ/112/742 J/ApJS/94/749 Hesselroth T.: J/AJ/131/250 Hessels J.W.T.: J/MNRAS/359/1524 Hesser J.: J/AJ/112/2408 Hesser J.E.: II/36 II/43 II/44 J/AJ/107/2067 J/AJ/110/1699 J/AJ/113/1026 J/AJ/117/247 J/AJ/119/1268

J/AJ/128/712 J/AJ/97/1360 J/ApJ/518/262 J/ApJ/527/199 J/ApJ/559/812 J/ApJS/120/265 J/PASP/109/883

J/PASP/110/1172 J/PASP/110/533 J/PASP/99/739 Hester J.: J/AJ/113/1001 J/PASP/108/682 Hester J.J.: J/AJ/104/92 J/AJ/112/628 J/AJ/117/1700 J/AJ/117/2757 Hetterscheidt M.: J/A+A/441/905 J/A+A/462/875 Heusler A.: J/A+A/390/39 J/A+A/406/L9 J/A+A/410/813 Hewett: J/MNRAS/325/1002 Hewett P.: J/MNRAS/224/453 J/MNRAS/345/1 J/MNRAS/351/1071 J/MNRAS/372/425 Hewett P.C.: J/AJ/109/1498 J/AJ/122/518 J/ApJ/445/62 J/ApJ/450/51 J/ApJS/134/35 J/MNRAS/352/285

VII/148 VII/182 VII/182 Hewish A.: J/MNRAS/229/589 HEWITT A.: J/ApJS/77/203 VII/105 Hewitt A.: VII/117 VII/158 VII/178 HEWITT A.: VII/27 VII/37 Hewitt J. N.: J/ApJ/510/64 Hewitt J.N.: VIII/52B Hewitt J.W.: J/ApJS/155/421 Heydari-Malayeri M.: J/A+A/347/841 J/A+A/381/941 J/A+A/422/129 J/A+A/436/117 Heyer I.: J/A+AS/130/245 J/AJ/112/1335 J/AJ/120/2747 Heyer M.H.: J/ApJ/551/852 HEYNDERICKX D.: J/A+AS/106/79 H.-G. and Gurtler: J/A+AS/120/429 Hibbard J.E.: J/AJ/127/660 J/ApJ/630/784 Hibbert A.: J/A+A/446/1185 Hickel G.: J/A+A/407/31 Hicken M.: J/AJ/130/2453 J/AJ/131/1648 J/AJ/131/527 Hickison P.: J/A+AS/94/327 Hicks E.K.: J/ApJS/142/79

Hickson P.: J/A+AS/110/313 J/A+AS/125/459 J/A+AS/146/373 VII/213 VII/213 VII/213 VII/213

VII/213 VII/85A Hidayat B.: J/A+AS/136/293 Hiesgen M.: J/A+AS/124/157 Higdon J.: J/AJ/130/1358 Higdon J.L.: J/ApJ/640/768 Higdon S.J.: J/ApJ/640/768 Higdon S.J.U.: J/ApJ/634/L1 Higgs L.A.: J/A+A/241/551 J/AJ/122/3155 J/ApJ/626/195 Higuera G.M.A.: J/PASP/114/233 HILAIRE G.: I/74 Hildebrand J.L.: J/ApJ/525/420 Hildebrand R.H.: J/ApJ/604/717 Hildebrandt G.: J/A+A/367/236 Hildebrandt H.: J/A+A/441/905 Hilditch R.W.: J/ApJS/115/185 Hilker M.: J/A+A/383/823 J/A+A/418/445 J/A+A/448/171 J/A+A/459/391 J/A+A/463/503 J/A+A/472/111

J/A+AS/134/59 J/A+AS/134/75 J/AJ/127/2114 J/AJ/127/980 Hill A.: J/AJ/131/414 Hill A.B.: J/ApJ/607/L33 J/ApJ/636/765 J/ApJS/170/175 Hill G.: J/A+A/324/965 J/A+A/328/551 Hill G.C.: J/other/APh/26.282 Hill G.J.: J/MNRAS/344/283 J/MNRAS/357/1231 Hill G.M.: J/A+A/276/142 J/A+A/335/973 J/A+A/361/991 J/A+A/458/569 Hill J.: J/AJ/126/119 Hill: J/A+A/475/915 Hill J.E.: J/A+A/448/L9 Hill J.K.: J/AJ/113/1011 J/AJ/116/180 HILL J.K.: J/ApJ/413/604 J/ApJ/425/122 Hill J.K.: J/ApJ/438/181 J/ApJS/98/595 HILL J.M.: J/AJ/104/2078 J/AJ/106/831 Hill J.M.: J/AJ/116/1529 J/AJ/120/2269 J/AJ/123/3018 J/AJ/124/1918 J/AJ/125/2393 J/AJ/131/2426

J/ApJ/416/36 J/ApJ/613/841 J/ApJS/115/1 J/ApJS/146/267 Hill R.: J/ApJ/463/26 Hill R.J.: J/ApJ/508/491 J/ApJ/521/155 J/ApJS/82/117 HILL R.J.: J/ApJS/91/583 Hill R.S.: J/AJ/112/601 J/ApJ/446/622 J/ApJ/548/127 J/ApJS/146/209 J/ApJS/148/97 HILL R.S.: V/51 Hill T.: J/MNRAS/363/405 J/MNRAS/368/1223 HILL V.: J/A+A/293/347 Hill V.: J/A+A/324/435 J/A+A/345/430 J/A+A/403/1105 J/A+A/414/503 J/A+A/416/1117 J/A+A/435/657

J/A+A/439/129 J/A+A/459/423 J/A+A/465/799 J/AJ/128/1177 HILLDRUP K.C.: VIII/29 Hillebrandt W.: J/MNRAS/369/1880 Hillenbrand L.: J/AJ/127/2659 J/AJ/128/787 Hillenbrand L.A.: J/AJ/106/1906 J/AJ/113/1733 J/AJ/114/198 J/AJ/116/1816 J/AJ/119/3026 J/AJ/120/1426

J/AJ/121/3160 J/AJ/123/1528 J/AJ/124/1001 J/ApJ/540/236 J/ApJ/574/258 J/ApJ/576/880 J/ApJ/584/911 J/ApJ/601/979 J/ApJS/160/319 J/PASP/118/1690

Hiller M.E.: J/A+A/371/79 Hills J.: J/AJ/119/1901 Hillwig T.C.: J/ApJ/639/1069 Hiltner W.A.: J/ApJS/2/389 J/ApJS/35/111 Hilton E.J.: J/AJ/121/2331 Hilton J.: J/A+AS/113/325 J/A+AS/126/73 Hilton J.L.: J/AJ/112/2319 Hily-Blant P.: J/A+A/469/173 Hindsley R.: I/270 I/273A J/AJ/126/1023 Hindsley R.B.: J/AJ/121/2308 J/AJ/121/2331 J/AJ/121/31 J/AJ/123/567 J/AJ/127/3034 Hines D.: J/ApJS/154/155 J/PASP/118/1690 Hines D.C.: J/ApJ/512/125 J/ApJ/569/23 J/ApJ/620/1010 Hinkle K.: III/240 J/ApJS/103/235 J/ApJS/107/312 J/PASP/107/1042 Hinshaw G.: J/ApJS/148/97 Hintz E.G.: J/AJ/132/393 HINTZEN P.: J/ApJ/413/604 Hintzen P.: J/ApJ/416/36 Hintzen P.M.N.: J/AJ/108/1350 J/AJ/112/601 J/ApJ/438/181 J/ApJS/98/595 Hippelein H.: J/A+A/285/785 J/A+AS/121/507 Hirabayashi H.: J/ApJ/616/110 J/ApJS/131/95 J/ApJS/155/33 J/PASJ/52/997 Hirao T.: J/AJ/117/446 Hirata R.: VI/69 Hirsch R.: J/A+A/465/331 Hirshfeld A.C.: J/A+A/313/841 Hirst P.: J/MNRAS/379/1343 Hirte S.: J/A+A/332/575 Hjalmarson A.: J/A+A/358/257 J/A+A/473/177 Hjalmarson A: J/A+AS/144/441 Hjalmarson A.: J/ApJS/117/427 Hjorth J.: J/A+A/391/481 J/A+A/455/L1 Ho L. C.: J/ApJ/510/659 J/ApJS/107/215

Ho L.C.: J/AJ/112/2408 J/AJ/121/3048 J/AJ/127/105 J/AJ/129/2138 J/ApJ/632/799 J/ApJ/641/L21

J/ApJS/112/315 J/ApJS/133/77 J/ApJS/98/477 Ho P.T.P.: J/A+A/474/911 Hoard D.W.: J/ApJ/565/511 J/ApJ/586/1356 Hoare M.G.: J/A+A/461/11 J/A+A/474/891 J/MNRAS/366/1096 J/MNRAS/379/663 Hobart M.: J/A+A/438/201 Hobart M.A.: J/A+AS/129/9 Hobbs G.: J/MNRAS/335/275 J/MNRAS/342/1299 J/MNRAS/352/1439 J/MNRAS/353/1311 J/MNRAS/360/974

J/MNRAS/362/1189 J/MNRAS/372/777 Hobbs G.B.: J/AJ/129/1993 Hobbs L.M.: J/ApJ/602/L117 J/ApJS/147/61 Hobbs M.: J/AJ/129/1993 Hobson M.P.: J/MNRAS/341/1066 HODAPP K.-W.: J/ApJS/94/615 HODDER P.J.C.: J/AJ/103/460 Hodge P.: J/ApJ/324/172 J/ApJS/73/661 J/PASP/97/753 Hodge P.W.: J/A+A/357/437 J/A+A/396/473 J/A+A/398/479 J/AJ/131/2478 J/ApJ/521/577 J/ApJ/559/851

J/ApJ/562/303 J/ApJ/610/247 J/MNRAS/300/665 J/PASP/111/685 Hodge T.M.: J/ApJS/151/299 Hodges J.: J/other/APh/26.282 Hodgkin S.: J/MNRAS/333/423 J/MNRAS/358/333 J/MNRAS/370/954 J/MNRAS/377/741 Hodgkin S.T.: J/A+A/448/579 J/A+A/463/1017 J/MNRAS/274/1194 J/MNRAS/313/347 J/MNRAS/372/1117

J/MNRAS/374/372 J/MNRAS/379/647 J/MNRAS/379/816 J/MNRAS/380/712 J/other/Sci/292.698 Hoeg E.: I/207 I/275 Hoeglund B.: J/A+AS/115/439 Hoeksema J.T.: J/A+A/390/717 Hoekstra H.: J/A+A/319/757 Hoekzemal N.M.: J/A+AS/98/505 Hoessel J. G.: J/PASP/108/682 HOESSEL J.G.: J/AJ/101/1561 Hoessel J.G.: J/AJ/113/1001 J/AJ/116/584 J/AJ/117/1700 J/AJ/125/1261 J/ApJ/241/486 J/ApJ/264/337

J/ApJ/306/30 J/ApJ/463/26 J/ApJ/507/655 J/ApJ/508/491 J/ApJ/512/48 J/ApJ/514/614 J/ApJ/521/155 J/ApJ/523/540 J/ApJS/60/507

Hoffleit D.: I/174 J/other/IBVS/5298 HOFFLEIT D.: V/25 Hoffleit D.: V/36B V/50 Hoffleit E.D.: I/238A Hoffman G.L.: J/AJ/130/2613 J/ApJ/473/822 J/ApJS/103/331 Hoffman K.D.: J/other/APh/26.282 Hoffmann M.: J/A+AS/124/291 Hoffmann W. F.: J/ApJS/122/221 Hoffmann W.F.: II/56 Hoflich P.: J/AJ/131/1648 Hofmann K.-H.: J/A+A/385/87 J/A+A/448/1075 Hofmann R.: J/ApJ/466/254 Hofmann W.: J/A+A/390/39 J/A+A/406/L9 J/A+A/410/813 Hofmeister A.M.: J/A+AS/146/437 J/MNRAS/371/1744 Hofner P.: J/A+AS/120/283 J/ApJS/154/553 J/ApJS/155/149 Hog E.: I/250 Hog E: I/259 Hog E.: I/276 HOG E.: I/62C I/97 Hogan C.: J/ApJ/594/1 Hogan E.: J/ApJ/612/511 Hogen R.H.: J/AJ/116/941 Hogerheijde: J/A+A/475/915 Hogg D.W.: J/AJ/120/2190 J/AJ/123/567 J/AJ/125/1817 J/ApJ/471/L5 J/ApJ/538/29 J/ApJS/127/1 Hogg H.S.: J/AJ/122/2587 Holberg J.B.: J/ApJ/571/512 J/ApJS/119/207 J/ApJS/147/145 J/ApJS/156/47 Holden B.: J/AJ/123/2223 J/AJ/128/1990 J/ApJ/621/53 Holden B.P.: J/AJ/120/1 J/AJ/127/2484 J/ApJ/644/30 J/ApJS/126/209 J/ApJS/142/153 Holdenried E.: J/AJ/127/539 Holdenried E.R.: I/268 I/289 J/AJ/119/3071 J/AJ/120/1120 J/AJ/121/3224 J/AJ/122/1586 J/AJ/124/2254 Holder B.P.: J/ApJ/581/325 HOLENSTEIN B.D.: J/A+AS/102/343 Holland S.: J/AJ/114/1488 Holland S.T.: J/ApJ/594/1 J/ApJ/602/571 Hollander A.: J/A+A/242/401 J/A+A/265/77 J/A+AS/101/87 Hollenbach D.: J/PASP/118/1690 Hollenbach D.J.: J/ApJ/633/871 Hollis J.: J/AJ/116/180 HOLLIS J.M.: VII/94 VIII/12 Holloway A.J.: J/A+A/472/805 Holm A.V.: II/83 Holman M.J.: J/AJ/131/527 J/ApJ/626/L5 Holmberg E.B.: VI/30A Holmberg J.: J/A+A/377/911 J/A+A/439/129 V/117 Holmes E.K.: J/ApJ/636/1098 Holmgren D.: J/A+A/319/867 J/A+A/367/236 Holmgren D.E.: J/A+A/405/1087 Holopainen J.: J/A+A/427/769 Holovatyy V.V.: J/AZh/79/867 J/AZh/82/437 Holt J.: J/MNRAS/348/857 Holt S.S.: J/ApJ/253/485 J/ApJS/96/303 Holtzman J. A.: J/PASP/108/682 Holtzman J.: J/AJ/113/1001 J/ApJ/633/844 Holtzman J.A.: J/AJ/117/1700 J/ApJ/448/179 Holtzman J.A: J/ApJ/459/L27 Holtzman J.A.: J/ApJ/478/124 J/ApJ/598/260 HOLWEGER H.: J/A+AS/42/391 Homan D.C.: J/AJ/130/1389 J/AJ/130/2473 J/AJ/131/1262 J/ApJ/609/539 Homeier D.: J/A+A/378/556 J/A+A/430/223 Homeier N.: J/A+A/391/857 J/A+A/435/65 J/AJ/128/1990 J/ApJ/621/53 Homeier N.L.: J/A+A/432/381 J/AJ/131/143 J/ApJ/644/30 Homer L.: J/AJ/126/2209 Honda M.: J/MNRAS/330/137 J/MNRAS/331/609 Honda S.: J/A+A/468/663 J/PASJ/57/13 J/PASJ/57/65 J/PASJ/58/389 Honeycutt R.: J/AJ/127/1158 Honeycutt R.K.: II/266 J/AJ/131/2673 J/ApJS/115/185 J/PASP/107/324 J/PASP/109/441 Hong K.S.: J/AJ/128/846 Hong X.-Y.: J/PASJ/52/997 Honma M.: J/ApJ/516/693 J/MNRAS/330/137 J/PASJ/52/631 Hony S.: J/A+A/384/473 J/MNRAS/376/313 Hoogerwerf R.: J/A+A/367/111 J/AJ/117/354 J/MNRAS/313/43 Hook I.: J/A+A/437/789 J/A+AS/137/75 Hook I.M.: J/A+A/386/97 J/A+A/399/469 J/AJ/112/407 J/AJ/127/2455 J/MNRAS/268/305 J/MNRAS/329/700 Hook R.: J/A+AS/137/75 J/A+AS/137/83 J/ApJ/613/200 Hook R.N.: J/AJ/112/1335 J/AJ/119/486 J/AJ/120/2747 HOOK R.N.: J/MNRAS/244/408 Hook R.N.: J/MNRAS/323/839 Hooper E.J.: J/ApJ/445/62 Hoopes C.: J/ApJS/144/21 Hopewell E.: J/MNRAS/362/689 Hopkins A.: J/MNRAS/305/297 J/MNRAS/306/708 J/PASJ/55/771 Hopkins A.M.: J/AJ/130/1358 J/ApJ/624/135 J/ApJS/155/1 J/MNRAS/296/839 Hopkins Ultraviolet Telescope Team: VI/103 Hopp U.: J/A+A/319/67 J/A+A/366/771 J/A+A/374/907 J/A+A/398/49 J/A+A/418/885 J/A+A/421/41

J/A+A/422/907 J/A+A/442/109 J/A+AS/109/537 HOPP U.: J/A+AS/111/527

Hopp U.: J/A+AS/116/43 J/A+AS/117/261 J/A+AS/134/317 J/A+AS/135/511 J/A+AS/142/399

J/A+AS/142/417 Hora J. L.: J/ApJS/122/221 Hora J.L.: J/ApJS/124/195 Horaguchi T.: VI/69 Horch E.: J/AJ/117/548 J/AJ/120/2638 J/AJ/122/1486 Horch E.P.: J/AJ/127/1727 J/AJ/131/1000 Horellou C.: J/A+AS/126/3 Horiai K.: J/PASJ/52/631 Horiuchi S.: J/ApJ/616/110 J/ApJS/155/33 J/PASJ/52/997 Hormaechea J.: J/A+AS/118/1 Horn J.: J/A+A/304/235 J/A+A/309/521 HORN J.: J/A+AS/107/403 Horne K.: J/A+A/365/110 J/A+A/369/57 J/A+A/379/564 J/A+A/467/785 J/ApJS/115/185 J/MNRAS/343/1131

J/MNRAS/358/795 J/MNRAS/362/117 J/MNRAS/372/1117 J/MNRAS/379/647 J/MNRAS/379/816 Horner D.J.: J/AJ/113/1788 J/ApJS/140/265 Hornoch K.: J/A+A/465/375 Horns D.: J/A+A/390/39 J/A+A/406/L9 J/A+A/410/813 Hornschemeier A.: J/ApJ/613/200

Hornschemeier A.E.: J/AJ/122/2810 J/AJ/124/2351 J/AJ/126/539 J/AJ/126/632 J/ApJ/554/742

J/ApJS/161/21 Hornstrup A.: J/ApJ/502/558 J/ApJ/594/154 Horst H.: J/A+A/473/369 Horvat A.: J/A+AS/125/399 Horvath J.E.: J/A+AS/136/261 J/A+AS/139/321 Hoshi T.: J/MNRAS/379/1546 Hoshina K.: J/other/APh/26.282 HOTTA N.: VII/140 Hou J.L.: J/AJ/125/1397 Houck J.R.: J/ApJ/634/L1 J/ApJ/635/280 J/ApJS/103/331 Houdashelt M.L.: J/AJ/119/1424 J/AJ/119/1448 J/ApJS/143/73 Houde M.: J/ApJ/604/717 Hough D.H.: J/AJ/130/23 Hough: J/A+A/385/1014 HOUK N.: III/133 Houk N.: III/214 III/31B III/51B III/80 Houziaux L.: II/86 HOUZIAUX L.: III/39A Houziaux L.: J/A+AS/131/469 HOUZIAUX L.: V/38 Hovatta T.: J/A+A/440/409 J/AJ/133/1947 Hoversten E.A.: J/AJ/127/3553 Hovhannesian R.Kh.: III/253 Hovhannisyan L.R.: J/A+A/449/425 Howard E.: II/246 Howard S.: VI/76 Howard T.A.: J/A+A/440/373 Howarth D.: J/MNRAS/375/145 Howarth I.D: J/A+A/315/95 Howarth I.D.: J/MNRAS/284/265 J/MNRAS/353/601 Howell D.A.: J/A+A/437/789 J/AJ/123/1411 Howell S.: J/AJ/112/2408 Howell S.B.: J/AJ/131/2673 J/MNRAS/369/1334 J/PASP/114/656 Howk J.C.: J/ApJ/617/718 J/PASP/118/1077 Howlett M.: J/AJ/124/690 Hoyos C.: J/MNRAS/365/454 Hoyt S.R.: J/AJ/121/2159 Hradecky V.: J/AJ/129/2856 Hric L.: J/A+A/373/950 J/A+A/448/1107 J/A+AS/135/29 Hron J.: J/A+A/338/581 J/A+A/406/51 HRON J.: J/A+AS/106/397 J/A+AS/99/31 Hrynek G.: J/A+AS/143/269 Hsieh B.C.: J/ApJS/158/161 HST Eta Carinae Treasury Project Team: J/A+A/419/215 Hu E.: J/ApJ/538/29 Hu E.M.: J/AJ/127/180 J/AJ/127/3121 J/AJ/127/3137 J/ApJ/471/L5 J/ApJS/94/461 Hu F.-X.: J/other/PPMtO/17b.1 Hu H.: J/A+A/383/1062 J/A+AS/118/239 Hu J.Y.: J/A+A/367/1049 J/A+AS/132/173 J/A+AS/139/575 J/AJ/130/1640 Hu Q.: J/A+A/466/627 Hu S.M.: J/MNRAS/373/209 Hua C.T.: J/A+A/268/419 J/A+A/309/446 Huang J.-S.: J/A+A/403/493 J/AJ/112/2471 J/ApJ/640/167 J/ApJ/642/126 J/ApJS/154/155 Huang J.S.: J/MNRAS/380/986 Huang L.: J/A+A/387/580 Huang M.: J/A+A/426/503 J/A+AS/115/283 J/ApJ/517/282 Huang W.: J/ApJ/639/1069 Huang X.-Y.: J/PASJ/52/997 Huang Z.: J/A+AS/125/475 Huang Z.-P.: J/ApJ/378/65 Huard T.L.: J/ApJ/638/293 Hube D.: J/A+AS/125/399 Hubeny I.: J/A+A/408/611 J/A+A/434/1125 J/A+A/436/687 J/A+A/466/277 J/ApJ/562/368 Huber M.E.: J/MNRAS/369/1334 Hubert A.-M.: J/A+A/304/235 J/A+A/328/551 Hubert A.M.: J/A+A/335/565 Hubert A.-M.: J/A+A/385/986 Hubert A.M.: J/A+A/409/275 J/A+A/411/565 Hubert A.-M.: J/A+A/440/305 J/A+A/451/1053 J/A+A/462/683 J/A+A/472/565 Hubert A.M.: J/A+A/472/577 Hubert D.: J/other/APh/26.282 Hubert H.: J/A+A/304/235 Hubrig S.: J/A+A/314/491 J/A+A/418/1073 J/A+A/425/263 J/A+AS/123/353 Hubscher J.: J/other/IBVS/5296 J/other/IBVS/5731 Huchra J.: II/246 IX/22 J/AJ/105/1637 J/AJ/116/102 J/ApJ/463/26 J/ApJ/507/655 HUCHRA J. P.: J/ApJ/384/404 Huchra J. P.: J/ApJ/510/659 Huchra J.: VII/75 HUCHRA J.P.: II/174 Huchra J.P.: J/A+A/369/57 J/A+A/371/79 J/AJ/106/1273 J/AJ/109/1458 J/AJ/109/874 J/AJ/112/1803

J/AJ/115/6 J/AJ/116/1 J/AJ/119/727 J/AJ/122/2458 J/AJ/122/2893 J/AJ/123/2490 J/AJ/131/1416 J/AJ/131/527 HUCHRA J.P.: J/ApJ/365/66 Huchra J.P.: J/ApJ/508/491 J/ApJ/512/48 J/ApJ/514/614 J/ApJ/521/155 J/ApJ/523/540 J/ApJ/529/723

J/ApJ/560/566 J/ApJ/617/192 J/ApJS/102/29 J/ApJS/115/185 J/ApJS/119/277 J/ApJS/121/287 J/ApJS/129/435 J/ApJS/149/29 J/ApJS/162/304 J/ApJS/99/391 J/MNRAS/274/1107 J/MNRAS/281/799 J/MNRAS/301/881 J/MNRAS/318/333 J/PASP/111/438

HUCHRA J.P.: VII/111 VII/129 VII/144 Huchra J.P.: VII/164 VII/185 VII/185 VII/185 VII/193 VII/213 VII/230 HUCHRA J.P.: VII/84A VII/86A Huchtmeier W.: J/A+AS/147/187

Huchtmeier W.K.: J/A+A/286/17 J/A+A/319/401 J/A+A/319/67 J/A+A/325/473 J/A+A/366/428 J/A+A/377/801 J/A+A/434/887 J/A+A/435/459 J/A+AS/139/97 J/A+AS/141/469 J/AJ/127/2031 J/PAZh/31/563 VIII/68

Huchtmeier W.K.H.: VIII/56 Hudec R.: J/A+A/235/185 HUDGINS D.M.: J/ApJS/86/713 Hudon J.D.: J/AJ/106/1059 HUDON J.D.: J/ApJS/85/315 Hudson C.E.: J/A+A/430/725 J/A+A/432/731 J/A+A/436/1131 J/A+A/452/1113 J/A+A/475/765 Hudson M.J.: J/AJ/128/1558 J/MNRAS/313/469 J/MNRAS/327/1004 J/MNRAS/327/249 J/MNRAS/327/265

J/PASP/111/57 Huebner W.F.: J/A+AS/120/301 Huelamo N.: J/A+A/359/227 J/A+A/460/133 Huenemoerder D.: J/ApJ/549/441 Huensch M.: J/A+A/402/571 J/A+AS/127/251 J/A+AS/132/155 J/A+AS/135/319 Huerta L.: J/ApJS/162/1 Huertas-Company M.: J/A+A/468/937 Huestamendia G.: J/A+A/307/80 Huetsi G.: J/A+A/459/L1 Huggins P.J.: J/ApJS/160/272 HUGGINS P.J.: VII/50 Hughes D.: J/MNRAS/344/887 Hughes D.H.: J/MNRAS/372/1621 Hughes H.S.: I/98A Hughes J.: J/AJ/112/2408 J/AJ/127/980 Hughes J. Wallerstein G.: J/AJ/119/1225 Hughes J.A.: I/100A Hughes S.: J/ApJS/154/623 Hughes S.M.: J/ApJ/463/26 J/ApJ/507/655 J/ApJ/523/540 Hughes S.M.G.: J/A+AS/145/11 J/AJ/99/784 J/ApJ/508/491 J/ApJ/512/48 J/ApJ/514/614 J/ApJ/521/155

J/ApJ/529/723 Hughey B.: J/other/APh/26.282 Huguenin D.: J/A+A/268/419 J/other/Ap+SS/163.275 Hui Bon Hoa A.: J/A+A/475/1053 Hui X.: J/ApJ/449/592 J/ApJ/458/455 J/ApJS/88/423 Hui X.-H.: J/ApJ/414/463 Huizinga J.E.: J/MNRAS/343/665 Huldtgren M.: J/A+A/350/883 J/A+A/357/219 J/A+A/372/173 J/A+A/421/623 Hulleman F.: J/A+A/416/1037 J/ApJ/563/L49 Hulme S.C.: J/AJ/128/712 Hulth P.O.: J/other/APh/26.282 Hultqvist K.: J/other/APh/26.282 Humason M.L.: J/ApJ/81/187 Hummel C.A.: J/AJ/114/2822 Hummel E.: J/A+AS/41/151 Hummel W.: J/A+A/312/818 Hummels C.B.: J/AJ/128/245 Hummer D.G.: VI/64 Humphrey P.J.: J/ApJ/612/848 J/MNRAS/344/134 Humphreys R.M.: II/144 II/145 II/254 HUMPHREYS R.M.: J/AJ/102/395 Humphreys R.M.: J/AJ/127/1567 J/AJ/96/1884 HUMPHREYS R.M.: J/PASP/104/553 V/44 Hundertmark S.: J/other/APh/26.282 Hunley J.: J/ApJ/510/659 Hunstead R.: J/A+A/392/795 Hunstead R.W.: J/AJ/123/1216 J/AJ/130/896 J/AJ/130/957 J/AJ/131/100 J/AJ/131/114 J/AJ/131/1280

J/MNRAS/285/898 J/MNRAS/324/1131 J/MNRAS/347/837 VIII/70A VIII/78 VIII/81 VIII/82 HUNT A.: J/MNRAS/251/330 Hunt A.J.: VIII/38 Hunt L. K.: J/ApJ/510/637 J/ApJS/125/349

Hunt L.K.: J/A+A/414/905 J/A+AS/120/521 J/A+AS/137/101 J/A+AS/142/65 J/ApJ/616/707

J/ApJ/626/L15 Hunt-Cunningham M.: J/MNRAS/363/405 Hunter D.A. Christian C.A.: III/92 Hunter D.A.: J/AJ/105/1196 J/AJ/113/1691 J/AJ/116/1275 J/AJ/116/146 J/AJ/128/2170 J/ApJ/448/179

J/ApJ/456/174 J/ApJ/459/L27 J/ApJ/468/633 J/ApJ/478/124 J/ApJS/162/49 J/ApJS/99/551 Hunter H.A.: J/ApJ/452/238 Hunter I.: J/A+A/436/687 J/A+A/437/467 J/A+A/457/265 J/A+A/466/277 J/A+A/471/625 J/MNRAS/367/1478 Hunter S.D.: IX/9 J/ApJS/101/259 J/ApJS/107/227 J/ApJS/123/79 J/ApJS/94/551 Hunter T.R.: J/A+A/466/989 J/ApJ/606/929 Hure J.M.: J/A+AS/117/561 Hurley J.R.: J/A+A/377/911 Hurley K.: J/A+A/438/1175 J/ApJ/615/887 J/ApJS/120/399 J/ApJS/156/217 J/other/GCN/1070.1 Hurley K.J.: J/A+A/385/377 Hurley-Keller D.: J/AJ/115/1840 J/AJ/115/1856 J/ApJ/616/804 Hurta Zs.: J/A+A/430/1049 J/A+A/445/617 Hurwitz M.: J/A+A/321/531 Hussain G.A.J.: J/MNRAS/284/265 Hutchings J.B.: J/AJ/116/941 J/AJ/126/63 J/ApJ/427/125 J/ApJ/469/629 J/ApJ/471/694 Hutchinson M.G.: J/A+A/412/405 HUTH H. and WENZEL W.: VI/28 Huth H.: VI/35 VI/38 VI/67

Hutsemekers D.: J/A+A/332/410 J/A+A/340/371 J/A+A/427/107 J/A+A/433/757 J/A+A/441/915

J/A+A/478/321 J/A+AS/142/451 Hutsi G.: J/A+A/462/811 Hutter D.J.: J/AJ/111/393 J/AJ/114/2822 Huub Rottgering: VIII/69A Huynh M.T.: J/AJ/130/1358 J/AJ/130/1373 Hwang H.S.: J/MNRAS/375/115 Hwang N.: J/A+A/454/717 Hyakkai M.: J/PASJ/56/295 Hyland A.R.: J/ApJS/69/99 J/MNRAS/291/261 J/MNRAS/301/640 Hyman S.D.: J/AJ/128/1646 Hyung S.: J/ApJ/614/745 J/PASP/105/1279 J/PASP/106/745

i Albertson C.: II/230 I.: J/A+A/465/271 J/MNRAS/273/1097 IANNA P.A.: I/212 Ianna P.A.: J/AJ/115/1076 Iannicola G.: J/A+A/330/453 J/AJ/121/916 I.A.U. Working Group on Reference Frames: I/218 Ibanez M.: J/A+AS/24/35 IBANOGLU C.: J/A+A/281/811

Ibanoglu C.: J/A+A/336/920 J/A+A/374/1056 J/A+A/388/298 J/A+A/405/1087 J/A+A/409/959

J/A+A/420/595 J/A+AS/129/9 J/MNRAS/370/2013 J/MNRAS/373/435 J/MNRAS/376/573 IBANOGLU M.C.: J/A+A/287/817 Ibata R.: J/A+A/353/958 J/A+A/440/1061 J/AJ/124/915 Ibata R.A.: J/AJ/113/634 J/AJ/127/2771 J/ApJS/120/265 IBATA R.A.: J/MNRAS/275/591 Ibata R.A.: J/MNRAS/339/701 J/MNRAS/375/831 J/MNRAS/380/281 Ibragimov M.A.: J/A+A/350/513 J/A+AS/139/513 J/PAZh/25/115 Ibrahimov M.A.: J/A+A/399/105 J/A+A/402/151 J/A+A/409/169 J/A+A/463/1017 J/MNRAS/310/210 Ibukiyama A.: J/A+A/394/927 Ichihashi K.: J/ApJ/382/636 Ichikawa R.: J/PASJ/52/631 Ichikawa S.: VII/147 Ichikawa T.: J/PASJ/33/107 Idiart T.P.: J/A+A/383/30 J/A+A/472/101 Idzi R.: J/AJ/124/2164

Iglesias-Paramo J.: J/A+A/384/383 J/A+A/397/421 J/A+A/401/471 J/A+A/419/47 J/A+A/425/429

J/ApJ/619/L95 J/ApJS/164/38 IGNJATOVIC LJ.: J/A+AS/103/57 Iijima T.: J/A+A/338/1006 Ikeda Y.: J/A+A/367/199 Ikhsanov N.R.: J/A+A/394/617

Ikonnikova N.P.: J/PAZh/25/30 J/PAZh/26/705 J/PAZh/27/187 J/PAZh/27/841 J/PAZh/29/858

J/PAZh/30/297 J/PAZh/30/855 J/PAZh/32/48 J/PAZh/32/662 J/PAZh/32/737 J/PAZh/33/679 Ikuta C.: J/MNRAS/366/144 Ilbert O.: J/A+A/403/857 J/A+A/428/1043 J/A+A/457/79 Iliadis C.: J/A+A/466/641 Iliev I.K.: J/A+A/443/157 J/A+A/454/171 J/A+A/459/871 J/A+A/462/591 Iliev I.Kh.: J/A+A/395/885 Iliev L.: J/A+A/395/885 Illiev L.: J/A+A/387/580 Illingworth G.: J/AJ/130/1 J/ApJ/459/110 J/ApJ/571/136 J/ApJ/621/53 J/ApJS/150/1 ILLINGWORTH G.D.: J/AJ/100/1091 Illingworth G.D.: J/AJ/128/1990 J/AJ/131/143 J/AJ/131/1891 J/ApJ/463/26 J/ApJ/481/673 J/ApJ/500/714

J/ApJ/507/655 J/ApJ/508/491 J/ApJ/512/48 J/ApJ/514/614 J/ApJ/521/155 J/ApJ/523/540 J/ApJ/529/723 J/ApJ/606/L25 J/ApJ/612/202 J/ApJ/624/L5 J/ApJ/644/30 J/ApJ/661/750 J/ApJS/142/1 J/ApJS/159/41

Ilovaisky S.A.: VII/15 Ilyin I.: J/A+A/334/863 J/A+A/360/499 J/A+A/380/578 J/A+A/402/745 J/A+A/431/933 J/A+A/433/185 Ilyin I.V.: J/A+A/341/553 J/A+A/476/881 J/A+AS/123/209 Im M.: J/AJ/131/250 J/ApJ/571/136 J/ApJS/142/1 J/ApJS/157/175 J/ApJS/159/41 Imai H.: J/ApJ/640/982 J/PASJ/52/631 J/PASJ/56/261 J/PASJ/56/765 Imanishi K.: J/PASJ/55/161 J/PASJ/55/635 J/PASJ/55/653 IMBERT M.: III/105 Imbert M.: J/A+A/387/850 J/A+AS/116/497 J/A+AS/140/79 Immler S.: J/ApJ/554/202 IMPEY C.: J/AJ/105/853 IMPEY C.D.: J/AJ/106/530 Impey C.D.: J/AJ/128/1974 J/AJ/128/544 J/ApJ/445/62 J/ApJ/524/536 J/ApJ/644/100 J/ApJS/105/209

J/ApJS/126/133 J/ApJS/128/99 J/ApJS/135/227 I.N.: J/PASP/109/849 in ’t Zand J.J.M.: J/A+A/472/705 Inada N.: J/AJ/131/1 J/MNRAS/382/412 Indebetouw R.: J/ApJ/635/560

Infante L.: J/A+A/383/823 J/A+A/418/445 J/A+A/463/503 J/A+A/472/111 J/A+AS/119/329 J/A+AS/120/357 J/A+AS/125/247 J/A+AS/134/59 J/A+AS/134/75 J/A+AS/134/75 J/AJ/121/148 J/AJ/127/1811 J/AJ/127/2114 J/AJ/128/1990 J/AJ/129/1237 J/AJ/130/2012 J/AJ/131/2035 J/AJ/134/1103 J/ApJ/612/202 J/ApJ/621/53 J/ApJS/126/1 J/ApJS/150/1 J/ApJS/162/1

INFANTE L.: VII/205 Infante L.: VII/208 Infrared Astronomical Satellite (IRAS) Catalogs and Atlases: II/274 II/275 Ingalls J.G.: J/ApJ/517/282 Ingerson T.: J/A+A/445/503 Inoue A.K.: J/ApJ/609/935 Inoue H.: IX/19 J/ApJ/518/656 Inoue M.: J/A+AS/39/379 J/ApJ/615/702 J/PASJ/52/631 J/PASJ/52/997 Institute of Astronomy: I/304 in’t Zand J.: J/A+A/455/813 Inui T.: J/PASJ/57/135 Ioka K.: J/ApJ/609/935 Iovino A.: J/A+A/375/1 J/A+A/403/857 J/A+A/428/1043 J/A+A/429/101 J/A+A/445/857 J/A+A/457/79

J/A+A/463/519 J/A+A/473/399 IOVINO A.: J/A+AS/112/347 Iovino A.: J/A+AS/119/265 J/AJ/125/1660 J/AJ/130/425 IPAC: VII/113 IPAC/UMass: II/241 VII/233 IPATOV A.: J/A+AS/112/439 Ipatov A.: J/A+AS/119/191 J/A+AS/125/1 J/AJ/112/2004 J/AJ/112/2020 J/AJ/114/189 J/AJ/116/2415 Ipatov A.P.: J/PAZh/21/596 I.P.Williams: J/A+AS/136/257 Irby B.K.: J/ApJS/158/139 Iriarte B.: II/5A Irsmambetova T.R.: J/A+AS/139/47 J/PAZh/21/652 J/PAZh/23/341 Irvine W.M.: J/ApJS/117/427 Irwin A.W.: III/185 Irwin J.: J/A+A/467/785 J/MNRAS/370/954 J/MNRAS/372/1117 J/MNRAS/377/741 J/MNRAS/379/647

J/MNRAS/379/816 Irwin J.A.: J/ApJ/598/969 J/ApJ/599/218 J/ApJ/600/729 J/ApJ/636/200 J/MNRAS/371/147 Irwin M.: J/A+A/353/958 J/A+A/390/491 J/A+A/392/851 J/A+A/440/1061 J/AJ/128/1177 J/AJ/130/2140

J/MNRAS/358/333 J/MNRAS/370/954 J/MNRAS/373/705 J/MNRAS/380/281 Irwin M.J.: I/282 J/A+A/437/467 J/A+A/457/265 J/A+A/459/423 J/A+AS/122/463 IRWIN M.J.: J/A+AS/91/171 Irwin M.J.: J/AJ/113/634 J/AJ/123/2490 J/ApJS/105/209 J/MNRAS/259/563 IRWIN M.J: J/MNRAS/261/657 Irwin M.J.: J/MNRAS/268/305 J/MNRAS/327/918 J/MNRAS/339/701 J/MNRAS/346/1103 J/MNRAS/353/601

J/MNRAS/369/97 J/MNRAS/375/831 VII/182 Isaev E.A.: VIII/67 ISENSEE J.E.: J/ApJ/425/122 Isensee J.E.: J/ApJS/98/595 Isern J.: J/A+A/465/249 J/A+AS/102/361 Ishida K.: J/MNRAS/238/709 Ishihara A.: J/MNRAS/370/1979 J/other/APh/26.282 Ishii H.: I/249 Ishii M.: J/AJ/117/446 J/ApJ/596/1064 Ishisaki Y.: J/ApJ/518/656 J/ApJS/133/1 J/ApJS/161/185 J/PASJ/53/445 Ishwara Chandra C.H.: J/A+A/472/557 Islas-Islas J.M.: J/AJ/129/220 ISO Data Centre: VI/111 Isobe N.: J/ApJ/645/256 Isobe S.: J/other/PNAOJ/2.561 Isobe T.: J/ApJ/589/509 ISRAEL F.P.: II/181 Israel F.P.: J/A+A/276/25

Israelian G.: J/A+A/398/363 J/A+A/415/1153 J/A+A/438/251 J/A+A/449/723 J/A+A/458/873

J/A+A/469/783 Israelinan G.: J/A+A/458/997 Ita Y.: J/MNRAS/353/705 J/MNRAS/370/1979 J/PASJ/56/261 J/PASJ/56/765 J/PASJ/58/529 Ito Y.: J/A+AS/125/399 Itoh K.: J/ApJ/645/256 Itoh N.: J/A+A/417/827 J/ApJ/382/636 J/ApJS/102/411 J/ApJS/63/661 J/ApJS/74/291 Itoh Y.: J/AJ/131/1608 J/PASJ/58/L29 Ivanov G.: I/261 J/A+AS/137/305 Ivanov G.A.: I/306 IV/28A IV/30 Ivanov G.R.: J/ApJS/89/85 Ivanov M.: J/A+AS/116/1 Ivanov M.M.: J/A+A/347/583 Ivanov V.D.: J/A+A/425/489 J/A+A/466/501 J/A+A/474/121 J/AJ/119/2274 J/ApJS/151/387 Ivans I. I.: J/PASP/111/812 Ivans I.I.: J/ApJ/645/613 Ivezi Z.: J/AJ/121/2331

Ivezic Z.: J/A+A/423/755 J/A+A/444/643 J/AJ/121/2308 J/AJ/121/31 J/AJ/122/503 J/AJ/123/1807 J/AJ/123/3409 J/AJ/123/567 J/AJ/124/1651 J/AJ/125/1817 J/AJ/126/1023 J/AJ/126/1720 J/AJ/126/2125 J/AJ/126/2209 J/AJ/126/2579 J/AJ/127/3034 J/AJ/127/3553 J/AJ/127/899 J/AJ/128/1002 J/AJ/128/426 J/AJ/129/2542 J/ApJ/588/824 J/ApJ/644/829 J/ApJS/155/257 J/ApJS/163/80 J/MNRAS/362/9 J/PASJ/55/771

Ivison R.: J/MNRAS/372/741

Ivison R.J.: III/174 J/A+AS/146/407 J/ApJ/581/844 J/ApJ/614/671 J/ApJ/616/71 J/ApJ/622/772

J/MNRAS/330/1 J/MNRAS/344/887 J/MNRAS/346/1125 J/MNRAS/371/963 J/MNRAS/372/1621 J/MNRAS/379/1343 Iwadate K.: J/PASJ/52/631 Iwamuro F.: J/AJ/131/1608 J/PASJ/53/653 J/PASJ/57/165 Iwata I.: J/ApJ/634/861 IWATA T.: III/177 III/177 Iwata T.: J/PASJ/53/517 Iye M.: J/ApJ/634/861 J/ApJ/648/7 J/ApJS/137/279 J/PASJ/53/653 J/PASJ/57/165 J/PASJ/58/313 Iyengar K.V.K.: J/A+AS/133/201 Iyomoto N.: J/ApJ/645/256 Iyudin A.: J/A+AS/143/145 Izmailov I.S.: I/292 Izotov Y.: J/A+A/419/469

Izotov Y.I.: J/A+A/366/771 J/A+A/374/907 J/A+A/421/539 J/A+A/442/109 J/A+A/448/955 J/A+A/464/885 J/A+AS/142/417 J/AJ/121/169 J/AJ/121/66 J/AJ/123/1292 J/ApJ/511/639 J/ApJ/602/200 J/ApJS/161/240

Izotov Yu.I.: J/A+AS/131/295 J/A+AS/135/511 J/A+AS/137/299 Izotova I.Y.: J/other/KFNT/22.187 Izumiura H.: J/PASJ/56/261 J/PASJ/56/765 J/PASJ/57/109 J/PASJ/57/13 J/PASJ/58/529 Izzard R.G.: J/A+A/466/641 J. E.: J/A+AS/135/145 J.: J/A+A/395/753 J/A+A/400/533 J/A+A/431/523 J/A+A/475/915 J/A+AS/114/109 J/A+AS/120/429

J/MNRAS/326/959 J.A.: J/A+A/364/391 JAAKKOLA T.: J/ApJ/413/453 Jaaniste J.: J/A+A/336/35 J/A+A/459/L1 J/A+A/462/811 J/A+AS/123/119 Jablonka P.: J/A+A/427/397 J/A+A/444/365 J/A+A/459/423 J/A+A/474/763 J/MNRAS/371/1777 Jablonski F.: J/A+A/465/993 J/A+AS/128/561 Jablonski F.J.: J/A+AS/124/55 J/A+AS/131/119 JABLONSKI M.: J/A+A/300/323 Jackel B.J.: J/ApJS/145/213 Jackson C.: J/MNRAS/329/87 J/MNRAS/352/939

Jackson C.A.: J/A+A/386/97 J/A+A/399/469 J/AJ/130/1358 J/AJ/130/1373 J/MNRAS/329/227

J/MNRAS/333/100 J/MNRAS/350/1485 J/MNRAS/352/1245 J/MNRAS/354/305 J/MNRAS/371/898 Jackson E.S.: I/223 I/227 I/228 I/231 Jackson J.: I/116 JACKSON J.: I/5 Jackson J.: I/6 Jackson J.M.: J/ApJ/517/282 J/ApJ/582/756 J/ApJ/635/560 J/ApJ/639/227 Jackson N.: J/A+AS/134/401 J/ApJS/107/175 J/MNRAS/326/1455 J/MNRAS/376/371 J/MNRAS/382/412 Jackson N.J.: J/MNRAS/377/430 VIII/72 Jackson R.: J/A+A/463/589 Jackson T.: J/ApJS/149/365 Jackson Z.C.: J/MNRAS/369/97 Jacobs C.S.: J/AJ/127/3587 Jacobsen J.: J/other/APh/26.282 Jacobsen T.S.: J/A+A/439/387 Jacobson H.R.: J/AJ/129/2725 Jacoby G.: J/A+A/373/536

Jacoby G.H.: III/92 J/A+A/306/9 J/A+A/419/563 J/AJ/123/269 J/AJ/124/1464 J/AJ/127/318 J/AJ/129/2232 J/AJ/131/2478 J/AJ/99/1079 J/ApJ/339/53 J/ApJ/344/704 J/ApJ/344/715 J/ApJ/356/332 J/ApJ/365/471 J/ApJ/383/487 J/ApJ/414/454 J/ApJ/414/463 J/ApJ/416/62 J/ApJ/417/209 J/ApJ/458/455 J/ApJ/462/1 J/ApJ/479/231 J/ApJ/492/62 J/ApJ/577/31 J/ApJ/582/170 J/ApJ/614/167 J/ApJ/616/804 J/ApJS/145/65 J/ApJS/38/351 J/ApJS/88/423

Jaeger A.C.: J/A+A/423/983 Jaeger H.: J/A+A/388/704 Jaeger K.: J/A+A/398/49 J/A+A/418/885 J/A+A/421/41 J/A+A/421/59 J/A+A/422/907 Jaehnig K.: II/249 Jaehrling R.: I/210 I/264 I/286 Jaervinen S.P.: J/A+A/432/657 J/A+A/440/735 Jaffe W.: J/A+A/428/877 J/AJ/108/1598 Jager C.: J/A+A/449/583 Jahn K.: J/A+A/295/101 Jahoda K.: J/A+A/418/927 J/ApJS/96/303 Jahreiss H.: I/196 I/210 I/264 I/286 J/A+A/353/958 J/A+A/442/211 JAHREISS H.: V/35 V/70A Jain A.: II/208 J/A+A/317/328 JAIN S.K.: J/A+AS/111/399 J/A+AS/111/407 Jakimova N.N.: VI/123 Jakobsen P.: J/A+A/287/769 J/A+A/288/413 Jakobsson P.: J/A+A/455/L1 Jamar C.: II/86 JAMAR C.: III/39A V/38 James D.: J/MNRAS/362/117 James D.J: J/A+A/399/105 James D.J.: J/A+A/448/189 J/A+A/450/147 James G.: J/A+A/464/939 J/A+A/464/953 J/A+A/464/967 James P.A.: J/A+A/414/23 JAMESON R.F.: J/A+AS/100/607 J/A+AS/109/29 Jameson R.F.: J/MNRAS/313/347 J/MNRAS/342/1241 J/MNRAS/365/447 J/MNRAS/373/95 J/MNRAS/374/372

J/MNRAS/380/712 Jancart S.: J/A+A/442/365 J/A+A/464/377 J/AJ/129/1001 Janes K.: J/AJ/112/2408 VII/11 Janes K.A.: II/45 J/AJ/106/181 J/AJ/108/585 J/AJ/110/652 J/AJ/113/1723 J/AJ/124/1570 J/MNRAS/364/179

J/MNRAS/371/1301 Jangren A.: J/AJ/127/1943 J/AJ/130/2571 J/AJ/130/2584 J/AJ/130/496 Janik J.: J/A+A/408/611 Janknecht E.: J/A+A/391/L11 J/A+A/458/427 Jankov S.: J/A+A/334/863 J/A+A/409/275 J/A+A/411/565 Jannuzi B.T.: J/AJ/127/213 J/AJ/128/1 J/ApJ/618/123 J/ApJ/639/816 J/ApJ/641/140 J/ApJ/645/955

J/ApJS/161/9 J/MNRAS/367/1261 J/PASP/106/646 Janot-Pacheco E.: J/A+A/451/1053 Jansen D.J.: J/A+A/279/541 Jansen R.A.: J/AJ/127/2002 J/ApJ/630/784 J/ApJ/639/724 J/ApJS/126/271 J/ApJS/126/331 J/PASP/117/227 Janssen T.: J/A+A/439/433 JANULIS R.: II/130 Janulis R.: J/A+A/459/557 Jao W.-C.: J/AJ/129/413 J/AJ/130/1658 Japaridze G.S.: J/other/APh/26.282 Jardine D.: J/A+A/465/331 Jarosik N.: J/ApJS/148/97 Jarrett T.: II/246 J/AJ/125/2521 J/AJ/129/2074 J/ApJS/127/1 Jarrett T.H.: J/A+A/405/99 J/A+A/408/465 J/A+A/408/67 J/AJ/128/1078 J/AJ/131/1416 J/ApJ/560/566

J/other/BSAO/57.5 J/PASP/112/1008 Jarvis: J/A+A/406/593 Jarvis M.: J/MNRAS/379/1343 Jarvis M.J.: J/AJ/123/1784 J/ApJS/150/417 J/MNRAS/355/20 J/MNRAS/372/1621 J/MNRAS/373/1531

J/MNRAS/375/1349 Jaschek C.: III/183 III/18B III/217 III/217 JASCHEK C.: III/53 Jaschek C.: III/83 J/A+AS/117/281 J/A+AS/118/495 J/A+AS/120/99 J/A+AS/123/135 J/A+AS/124/441

J/A+AS/128/475 J/A+AS/131/479 JASCHEK C.: V/25 V/61 Jaschek C.: VI/57 Jaschek M.: III/183 III/217 III/217 III/42 JASCHEK M.: III/53 Jaschek M.: III/67A JASCHEK M.: III/68 Jaschek M.: III/83 J/A+AS/117/281 J/A+AS/118/495 J/A+AS/120/99 J/A+AS/123/135 J/A+AS/124/441

VI/57 Jasinta D.M.D.: J/A+AS/107/235 J/A+AS/114/487 J/A+AS/118/381 J/A+AS/134/87 J/A+AS/136/293 Jasniewicz G.: II/75 JASNIEWICZ G.: J/A+A/287/591 Jasniewicz G.: J/A+A/423/353 Jauncey D.L.: J/AJ/110/880 J/AJ/127/3609 J/AJ/128/1570 J/AJ/130/2529 J/ApJ/614/607 J/MNRAS/325/1411

J/other/PASAu/10.310 J/PASJ/52/997 J/PASJ/55/351 VIII/34 VIII/61 JAUNSEN A.O.: J/A+A/300/323 Jaunsen A.O.: J/A+A/391/481 Javiel S.C.: J/A+A/458/113 Jayawardhana R.: J/AJ/131/527 J/ApJ/610/1064 J.D.: J/ApJS/119/189 JEANDENANS P.: I/74 Jedrzejewski R.: J/A+A/287/769 J/A+A/288/413 Jee M.J.: J/ApJ/644/30 Jeffers S.: J/A+A/475/1053 JEFFERY C.S.: J/A+AS/103/445 Jeffery E.J.: J/ApJS/156/227

Jeffries R.D.: J/A+A/375/863 J/ApJ/588/1009 J/ApJ/606/466 J/MNRAS/335/291 J/MNRAS/341/805 J/MNRAS/346/1143 J/MNRAS/351/1401 J/MNRAS/356/1583 J/MNRAS/356/89 J/MNRAS/358/13 J/MNRAS/358/341 J/MNRAS/364/272 J/MNRAS/367/781 J/MNRAS/375/1220 J/MNRAS/376/580 Jehin E.: J/MNRAS/367/1478 Jelinek M.: J/ApJ/641/L13 Jelinski P.N.: J/ApJS/164/38 Jelinsky P.: II/203A IX/35 J/AJ/107/751 J/ApJS/94/127 Jelinsky P.N.: J/AJ/130/825 J/ApJ/619/L95 Jenkins C.J.: VIII/32 Jenkins E.B.: III/77 J/ApJ/636/631 J/ApJS/146/1 J/ApJS/146/125 VIII/7A Jenkins J.S.: J/MNRAS/372/163 Jenkins L.F. (1963): I/60 Jenkins L.F.: I/10 JENKINS L.F.: I/81 Jenkins L.P.: J/MNRAS/357/401 Jenkins R.: J/PASJ/52/997 Jenkner H.: I/220 I/254 I/255 Jenness: J/A+A/385/1014 Jenness T.: J/MNRAS/344/887 J/MNRAS/372/1621 JENNINGS D.: J/MNRAS/251/330 Jennings M. C.: J/ApJ/446/115 Jenniskens P.: J/A+A/275/549 J/A+A/287/990 J/A+A/295/206 J/A+AS/128/179 Jensen H. R.: J/A+A/465/965 Jensen K.S.: III/75B JENSEN K.S.: J/A+AS/112/123 Jensen K.S.: J/A+AS/115/315 Jeon Y.-B.: J/AJ/126/2840 Jeong K.S.: J/A+A/361/641 Jerius D.: J/AJ/131/527

Jerjen H.: J/A+A/327/952 J/A+AS/124/1 J/AJ/123/873 J/AJ/124/1954 J/AJ/124/690 J/AJ/128/16

J/MNRAS/337/641 Jerjen: J/A+A/448/983 Jernigan J.G.: J/other/GCN/1070.1 JERZYKIEWICZ M.: J/MNRAS/260/782 Jester S.: J/AJ/126/2579 J/AJ/130/873 J/AJ/131/84 Jestu L.: J/A+A/276/345 JETSU L.: J/A+A/278/449 J/A+A/282/L9 Jetsu L.: J/A+A/309/655 J/A+A/321/L33 J/A+A/350/513 J/A+A/351/212 J/A+A/362/223 JETSU L.: J/A+AS/103/183 Jetsu L.: J/A+AS/139/513 Jetzer P.: J/MNRAS/351/1071 JEWELL P.R.: VIII/12 Jewell S.J.: J/MNRAS/274/1194 Jewitt D.: J/AJ/127/3023 J.G.: J/A+A/416/507 J.H.: J/A+A/385/1014

Jha S.: J/AJ/130/2453 J/AJ/131/1648 J/AJ/131/527 J/ApJ/594/1 J/ApJ/602/571 J/ApJ/607/665

J/ApJ/627/579 J.H.M.M.: J/A+A/403/247 Ji H.R.: J/A+AS/127/181 J/other/PPMtO/14.185 Ji Z.: J/A+AS/126/503 Jia Guibin: J/A+AS/114/337 J/A+AS/128/315 J/A+AS/129/569 Jiang B.W.: J/A+A/400/903 J/A+A/446/551 J/ApJS/106/463 Jiang D.R.: J/A+A/466/63 Jiang D.-R.: J/other/ChJAA/3.49 Jiang P.: I/258 J/A+A/474/783 Jiang P.F.: I/257B I/257B Jiang S.Y.: J/A+A/336/920 Jiang S.-Y.: J/A+AS/115/469 Jiang X.: J/A+A/459/557 Jiang Y.: J/A+A/312/879 Jiang Z.: J/A+A/397/361 J/A+A/449/143 J/AJ/117/2757 J/AJ/122/1718 J/AJ/126/1286 J/AJ/127/2579

J/AJ/129/1818 J/AJ/131/2245 J/ApJ/596/1064 J/ApJS/141/157 J/ApJS/149/53 Jiang Z.J.: J/AJ/112/628 Jiang Z.-J.: J/PASP/118/1104 Jijina J.: J/ApJS/125/161 Jilinski E.: J/A+A/448/1001

Jilinski E.G.: J/A+A/384/650 J/A+A/396/125 J/A+A/401/531 J/A+A/451/901 J/A+AS/117/335

J/A+AS/135/227 J/A+AS/143/265 Jimemez R.: J/A+A/317/54 Jimenez-Esteban F.M.: J/A+A/431/779 J/A+A/446/773 J/A+A/458/533 Jimenez-Reyes S.J.: J/A+A/390/717 Jimenez-Serra I.: J/A+A/469/207 Jimenez-Vicente J.: J/A+A/382/513 J/MNRAS/371/703 J/MNRAS/374/664 Jin H.: J/A+A/404/621 Jin W.: J/A+AS/120/201 J.J.: J/A+AS/139/29 Joblin C.: J/A+A/469/575 Jobse K.: J/A+AS/128/179 Jockers K.: J/A+A/383/724 J/other/KFNT/15.483 JODINSKIENE E.: II/157A Jodinskiene E.: II/93 IV/20 Jodrell Bank: B/merlin Joench-Soerensen H.: J/A+A/304/415 Joergensen H.E.: J/A+A/435/781 JOERGENSEN H.E.: J/A+AS/104/179 Joergensen Inger: J/PASP/106/967 Joeveer M.: J/A+A/459/L1 J/A+A/462/811 J/AN/327/365 Jog C.J.: J/A+A/438/507 Johansson L.E.B.: J/A+A/276/25 J/A+AS/115/439 J/AZh/79/610 Johansson P.: J/A+A/383/197 Johansson S.: J/A+A/377/361 J/A+A/419/215 JOHANSSON S.: J/A+AS/102/269 Johansson S.: J/ApJS/94/221 Johns-Krull C.M.: J/AJ/112/1197 J/ApJS/147/305 J/ApJS/165/256 Johnson D.: J/A+A/374/1056 Johnson H.: II/230 Johnson H.L.: II/52 II/5A II/84 II/85 JOHNSON H.L.: III/44 Johnson H.L.: III/45 Johnson J.: J/AJ/105/980 Johnson J.A.: J/AJ/117/247 J/AJ/128/2402 J/ApJ/527/199 J/ApJ/595/1154 Johnson K.E.: J/AJ/117/1708 J/ApJ/454/151 J/ApJ/559/864 Johnson N.: J/A+A/475/1053 Johnson O.: J/AJ/129/2074 J/ApJ/505/793 J/MNRAS/343/293 J/MNRAS/355/97 J/MNRAS/371/1777 Johnson R.A.: J/A+A/385/67 Johnson S.B.: J/A+AS/119/1 Johnston D.: J/AJ/128/1748 J/AJ/129/2007 J/AJ/130/832 Johnston D.E.: J/AJ/127/1811 J/AJ/131/1 Johnston H.M.: J/A+A/309/116 Johnston K.: J/A+A/459/557 JOHNSTON K.J.: I/151 I/169 Johnston K.J.: J/AJ/110/880 J/AJ/127/3609 J/AJ/128/1570 J/AJ/130/2529 J/ApJ/614/607 J/ApJS/136/265 Johnston K.V.: J/AJ/128/245 Johnston L.: II/88 Johnstone D.: J/ApJ/559/307 J/ApJ/639/259 Johnston-Hollitt M.: J/AJ/123/1216 J/AJ/130/957 J/AJ/131/1280 Joint IRAS Science W.G.: II/125 JOINT IRAS SCIENCE W.G.: VII/91 JOINT IRAS SCIENCE WORKING GROUP: II/102 Joint IRAS Science Working Group: III/197 Joly M.: J/A+A/417/515 J/A+A/475/487 J/A+AS/116/239 J/A+AS/117/93 Jomaron C.M.: J/MNRAS/274/1194 Jonauskas V.: J/A+A/416/383 J/A+A/424/363 J/A+A/433/745 J/A+A/455/1157 Joner M.D.: J/AJ/132/111 J/ApJS/159/100 Jones A.: J/A+AS/116/1 J/other/APh/26.282 VII/107 Jones A: VII/142A Jones B.: J/A+A/310/31 J/A+A/310/8 J/AJ/127/2659 J/AJ/128/787 J/MNRAS/350/1485 J/MNRAS/362/1143 Jones B.B.: J/ApJS/107/227 J/ApJS/123/79 Jones B.F.: I/199A I/200 I/214 I/283A I/293 JONES B.F.: J/AJ/105/2299 Jones B.F.: J/AJ/106/1059 JONES B.F.: J/AJ/106/1080 Jones B.F.: J/AJ/111/1193 J/AJ/123/1528 JONES B.F.: J/ApJS/85/315

Jones C.: IX/13 J/ApJ/412/479 J/ApJ/502/558 J/ApJ/511/65 J/ApJ/519/533 J/ApJ/560/675

J/ApJ/562/254 J/ApJ/594/154 J/ApJ/612/511 J/ApJ/635/1053 J/ApJS/111/163 J/ApJS/161/9 J/MNRAS/292/419 JONES C.: VII/18 Jones C.J.: J/ApJ/641/140 Jones D.H.: VII/249 Jones D.H.P.: J/A+AS/121/65 J/A+AS/130/77 JONES G.: J/ApJS/78/239 Jones H.R.A.: J/A+A/397/575 J/A+A/416/655 J/A+A/421/763 J/MNRAS/342/1241 J/MNRAS/370/163

J/MNRAS/372/163 VI/121 Jones J.B.: J/A+A/476/59 J/AJ/109/1095 Jones J.H.: J/AJ/105/1196 Jones L.: J/MNRAS/382/279 Jones L.R.: J/AJ/115/1253 J/ApJS/140/265 J/MNRAS/251/482 J/MNRAS/342/802 Jones M.E.: J/MNRAS/341/1066 J/MNRAS/342/915 J/MNRAS/360/340 Jones P.: J/MNRAS/367/1609

Jones P.A.: J/A+AS/111/311 J/A+AS/120/77 J/A+AS/121/321 J/A+AS/127/119 J/A+AS/130/421

J/ApJS/80/137 J/MNRAS/268/602 J/MNRAS/335/1085 J/MNRAS/355/44 J/other/SerAJ/166.13Short Jones T.E.: VI/105 JONES T.J.: J/AJ/107/1433 Jones T.J.: J/AJ/107/2120 J/AJ/110/268 J/AJ/96/1884 Jones T.W.: J/ApJ/613/343 Jonsell K.: J/A+A/440/321 Joo J.-M.: J/AJ/119/1824 J/AJ/127/958 Jopek T.J.: J/A+A/320/631 Jorda L.: J/A+AS/122/415 Jordan A.: J/A+A/472/111 J/AJ/127/24 J/ApJ/613/279 J/ApJS/153/223 J/ApJS/154/509 Jordan C.: J/MNRAS/353/1311 J/MNRAS/361/1102 Jordan C.A.: J/MNRAS/342/738 Jordi C.: J/A+A/324/549 J/A+A/326/1044 J/A+A/333/471 J/A+A/337/125 J/A+A/381/464 J/A+A/426/819

J/A+A/426/827 J/A+A/437/457 J/A+A/450/735 J/A+A/459/321 J/A+A/470/585 J/A+A/478/507 JORDI C.: J/A+AS/114/489 Jordi C.: J/A+AS/115/401 J/A+AS/128/265 J/A+AS/131/239 J/A+AS/146/169 JORDI C.: J/A+AS/94/519 Jore K.: J/A+AS/116/193 Jorge S.: J/A+A/458/997 Jorgensen A.M.: J/AJ/124/2164 Jorgensen B.R.: V/117 JORGENSEN H.E.: J/A+A/279/61 Jorgensen H.E.: J/A+A/388/1 J/A+A/409/439 J/A+A/442/841 J/A+A/456/881 Jorgensen I.: J/AJ/123/1411 J/AJ/127/2455 J/AJ/130/445 Jorgensen: J/A+A/475/915 Jorgensen J.K.: J/A+A/475/549 Jorgensen: J/MNRAS/273/1097 Jorgensen U.G.: J/A+A/317/54 JORISSEN A.: II/200

Jorissen A.: II/202 J/A+A/304/415 J/A+A/372/85 J/A+A/379/305 J/A+A/397/997 J/A+A/399/1167 J/A+A/403/993 J/A+A/430/165 J/A+A/441/1195 J/A+A/442/365 J/A+A/464/377 J/A+AS/113/51 J/A+AS/129/363

Jorsater S.: J/A+A/295/585 Jorstad S.G.: J/A+A/385/55 J/AJ/130/1418 Joseph J.M.: J/other/APh/26.282 Josephs T.S.: J/ApJ/525/420 Joshi B.C.: J/MNRAS/342/1299 J/MNRAS/352/1439 Joshi M.: J/MNRAS/351/1290 Joshi M.N.: J/A+AS/80/215 Joshi S.: J/A+A/455/303 J/MNRAS/376/371 Joshi Y.C.: J/A+A/402/113 Josselin E.: J/A+A/357/225 Jourdain de Muizon M.: J/A+A/365/545 Journet A.: J/A+AS/137/269 JOUTEL F.: VI/63 Joutsamo O.: J/A+A/427/769 J/A+AS/132/305 Joy A.H.: J/ApJ/81/187 J/ApJS/28/1 J.P.: J/A+A/394/395 J.P: J/MNRAS/325/1002 Ju B.: J/ApJS/141/157 Ju K.H.: J/A+AS/49/715 Juchert M.: J/A+A/447/921 Jugaku J.: J/MNRAS/316/326 J/MNRAS/330/137 Juhola M.: J/A+A/427/769 JULIEN P.: VII/18 Jung I.: J/A+A/390/39 J/A+A/406/L9 J/A+A/410/813 JUNG J.: I/83 IV/5 Jung J.: IV/6 Jungwiert B.: J/A+A/456/953 J/A+A/459/55 JUNKKARINEN V.: J/ApJS/77/203 JUNKKARINEN V.T.: J/AJ/106/426 Junqueira S.: J/A+A/333/658 Jura M.: J/AJ/116/941 JURA M.: J/ApJS/70/329 Juraszek S.: J/AJ/124/1954 J/AJ/128/16 Jurcsik J.: J/A+A/312/111 J/A+A/430/1049 J/A+A/445/617 J/A+A/469/1033 J/AcA/50/221 Jurdana-Sepic R.: J/A+A/370/503 J/A+A/386/237 JURRIENS T.A. KESTER D.J.M.: II/148 JURRIENS Th.: II/151 Juvela M.: J/A+AS/118/191 JUZA K.: J/A+AS/107/403 K.: J/A+A/374/504 J/A+A/374/504 J/A+A/454/855 J/A+A/475/915 Kaaret P.: J/ApJ/578/114

Kaas A.A.: J/A+A/350/883 J/A+A/357/219 J/A+A/372/173 J/A+A/391/481 J/A+A/420/945

J/A+A/421/623 Kaasalainen M.: J/A+A/465/331 Kaasalainen S.: J/A+A/374/1049 Kaastra J.S.: J/A+A/376/1113 J/A+A/420/135 J/A+AS/97/443 Kabe S.: J/MNRAS/330/137 Kablan H.: J/A+A/371/1078 Kadla Z.I.: II/248 Kadler M.: J/AJ/130/2473 J/ApJ/609/539 Kadota K.: II/273 Kaewkornmaugh P.: J/A+A/435/1099 Kafka S.: J/AJ/131/2673

Kahabka P.: J/A+A/305/74 J/A+A/361/491 J/A+A/388/100 J/A+AS/127/119 J/A+AS/136/81

J/A+AS/142/41 Kahabka P.T.: J/ApJ/448/683 Kahane C.: J/A+A/314/871 J/A+AS/142/181 Kahanpaa J.: J/A+A/350/513 Kahanpaeae J.: J/A+A/405/999 Kaiden M.: VII/244A Kaifu N.: J/AJ/131/1608 J/PASJ/57/165 Kainulainen J.: J/A+A/463/1029 Kaiser C.R.: J/ApJ/636/1002 Kaiser M.E.: J/AJ/119/486 J/AJ/127/1360 Kaiser N.: J/ApJ/602/571 J/ApJ/625/6 J/MNRAS/308/897 Kajisawa M.: J/ApJ/634/861 Kajsin S.: J/A+AS/124/559 Kalandadze N.B.: III/230 Kalberla P.M.W.: VIII/76 Kalberta P.M.W.: VIII/75 Kaldare R.: J/MNRAS/339/652 Kalenskii S.V.: J/A+AS/134/115 J/AZh/71/37 J/AZh/72/22 J/AZh/79/610 Kaler J.B.: J/A+AS/94/399 J/ApJ/417/209 J/PASP/109/289 Kalinichenko O.A.: I/292 J/PAZh/24/868 Kalinkov M.: J/A+AS/113/451 VII/190 VII/201 Kalirai J.: J/AJ/127/2771 Kallarakal V.V.: J/AJ/105/1571 Kallman T.R.: J/A+A/410/359 J/A+A/414/377 J/A+A/418/1171 Kalloghlian A.T.: J/other/Ap/41.1 Kallrath D. H. Bradstreet J.: J/AJ/109/359 Kallrath J.: J/A+A/362/673 J/A+A/452/959 J/AJ/110/2400 Kalomeni B.: J/A+A/405/1087

Kaltcheva N.T.: J/A+A/352/600 J/A+A/352/605 J/A+A/372/95 J/A+A/410/523 J/A+AS/128/309

J/A+AS/146/365 Kaluzny J.: I/244A J/A+A/295/101 J/A+A/373/899 J/A+A/424/1101 J/A+AS/108/151 J/A+AS/112/407

J/A+AS/114/1 J/A+AS/118/303 J/A+AS/120/139 J/A+AS/121/455 J/A+AS/122/1 J/A+AS/122/471 J/A+AS/125/337 J/A+AS/125/343 J/A+AS/128/19 J/A+AS/133/25 J/AcA/48/455 J/AJ/110/1141 J/AJ/110/2206 J/AJ/115/1016 J/AJ/115/1894 J/AJ/117/2810 J/AJ/118/2211 J/AJ/118/346 J/AJ/121/2032 J/AJ/121/861 J/AJ/121/870 J/AJ/122/1383 J/AJ/122/2477 J/AJ/124/1486 J/AJ/126/175 J/AJ/126/237 J/AJ/129/1596 J/AJ/131/407 J/MNRAS/243/492 J/MNRAS/365/110 J/MNRAS/371/1405

Kamaeva L.A.: J/AZh/81/209 Kamath U.S.: J/MNRAS/353/1025 Kambe E.: J/ApJ/525/420 J/PASJ/57/109 J/PASJ/57/13 J/PASJ/57/27 Kameno S.: J/PASJ/52/631 J/PASJ/52/997 KAMENO S.: VII/140 Kameya O.: J/PASJ/52/631 J/PASJ/56/261 J/PASJ/56/765 J/PASJ/58/529 Kaminker A.D.: J/A+A/412/707 Kaminski K.: J/A+A/465/331 Kaminskiy B.: VI/121 Kamohara R.: J/ApJ/640/982 J/PASJ/52/631 Kamperman T.M.: J/A+A/287/769 J/A+A/288/413 Kampert K.-H.: J/other/APh/26.282 Kamphuis J.J.: J/A+AS/116/15

Kanbach G.: J/ApJ/615/897 J/ApJS/101/259 J/ApJS/107/227 J/ApJS/120/409 J/ApJS/123/79

J/ApJS/94/551 Kanbur S.: J/A+A/411/361 Kanbur S.M: J/MNRAS/369/723 Kandalian R.A.: J/other/Ap/41.1 Kandori R.: VII/244A Kane S.: J/MNRAS/343/1131 Kane S.R.: J/A+A/467/785 J/MNRAS/362/117 J/MNRAS/372/1117 J/MNRAS/379/647 J/MNRAS/379/816 Kaneda H.: J/ApJ/635/214 J/ApJ/645/256 J/ApJS/134/77 Kaneko Y.: J/PASJ/52/631 Kanevsky B.: J/PASJ/52/997 Kang A.: J/A+A/437/61 J/A+A/445/69 J/A+A/454/717 Kang D.: J/AJ/127/539 Kang J.: J/ApJS/165/360 Kang Y.H.: J/MNRAS/310/982 Kang Y.W.: J/AJ/128/846 J/MNRAS/359/865 Kania U.: J/A+AS/120/267 Kankanyan R.: J/A+A/390/39 J/A+A/406/L9 J/A+A/410/813 Kannappan S.J.: J/AJ/131/527 Kantcen L.E.: J/AZh/86/158 Kan-Ya Y.: J/MNRAS/330/137 Kapahi V.K.: J/ApJS/107/19 KAPER L.: J/A+A/284/515 Kaper L.: J/A+A/377/925 J/A+A/384/473 J/A+A/428/149 J/A+A/429/559 J/A+A/430/137 J/A+A/464/581

J/A+A/473/523 J/A+AS/116/257 Kaplan D.L.: J/ApJS/117/361 J/ApJS/123/219 Kaplan G.H.: J/AJ/129/2420 Kaplan J.: J/A+A/421/509 J/MNRAS/351/1071 Kaplenko V.V.: J/A+AS/132/233 Kapteyn J.C.: I/108 Karaali S.: J/AN/326/321 J/MNRAS/355/307 Karachentsev I.: J/A+A/315/348

Karachentsev I.D.: J/A+A/296/643 J/A+A/366/428 J/A+A/377/801 J/A+A/396/473 J/A+A/398/479 J/A+A/433/751 J/A+A/435/459 J/A+AS/119/499 J/A+AS/127/409 J/A+AS/131/1 J/A+AS/133/181 J/A+AS/139/97 J/A+AS/141/469 J/A+AS/142/425 J/A+AS/147/187 J/AJ/114/122 J/AJ/127/1883 J/AJ/127/2031 J/AJ/130/2613 J/AN/322/73 J/AN/325/307 J/AZh/77/569

KARACHENTSEV I.D.: J/MNRAS/243/390 Karachentsev I.D.: J/other/BSAO/50.5 J/other/BSAO/57.5 J/PAZh/31/332 J/PAZh/31/563 VII/162 VII/219

VII/77 KARACHENTSEV I.D.: VII/83 Karachentseva V.E.: J/A+A/296/643 J/A+A/366/428 Karachentseva V.E: J/A+A/377/801 Karachentseva V.E.: J/A+A/396/473 J/A+A/398/479 J/A+A/435/459 J/A+AS/127/409 J/A+AS/141/469

J/A+AS/147/187 J/AJ/127/2031 J/AJ/130/2613 J/AZh/77/569 J/other/BSAO/50.5 J/other/BSAO/57.5 J/PAZh/31/563 VII/162 VII/219 VII/82A

KARACHENTSEVA V.E.: VII/83 Karakas A.I.: J/A+A/466/641 Karakashian T.: IX/13 Karakus O.: J/A+A/391/641 Karanchentsev I.: J/A+AS/124/559 J/A+AS/128/459 Karatas Y.: J/MNRAS/349/1069 J/MNRAS/357/497 J/MNRAS/360/1345 J/MNRAS/371/1793 J/MNRAS/373/1483 Karbovsky V.: J/A+A/438/377 Kardashev N.S.: J/AJ/130/2473 J/AZh/81/925 J/AZh/84/387 Karick A.M.: J/MNRAS/344/188 KARIMOVA D.K.: I/150 Karitskaya E.A.: II/140 II/205 J/AZh/78/408 J/PAZh/26/27 Karl C.: J/A+A/378/556 J/A+A/442/1023 Karlamaa K.: J/A+A/427/769 J/A+AS/116/157 J/A+AS/132/305 Karle A.: J/other/APh/26.282 Karle J.H.: J/AJ/106/2058 Karoji H.: J/MNRAS/316/326 J/PASJ/57/165 Karovska M.: J/ApJS/150/19 Karpischek R.U.: J/A+A/381/32 Karpov N.: J/A+A/383/724 Karpov N.V.: J/other/KFNT/15.483 Karpowicz M.: VII/4A Karttunen H.: J/A+A/374/1049 Kashikawa N.: J/ApJ/648/7 J/ApJS/137/279 J/MNRAS/379/1546 J/PASJ/57/165 J/PASJ/58/313 KASHIKAWA N.: VII/140 Kashyap V.: J/A+A/341/751 J/ApJ/644/829 J/ApJS/102/75 J/ApJS/150/19 Kasjan S.: J/A+A/438/377 Kasliwal M.M.: J/ApJ/634/L1 Kaspi S.: J/A+A/369/57 J/A+A/410/879 J/ApJ/510/659 J/ApJ/613/682 J/ApJ/632/799 J/ApJS/115/185 Kaspi V.M.: J/MNRAS/328/17 J/MNRAS/335/275 J/MNRAS/342/1299 J/MNRAS/352/1439 Kassim N.E.: J/A+A/456/791 J/AJ/128/1646 J/AJ/132/242 J/ApJ/591/640 J/ApJS/150/417 J/ApJS/68/715

J/MNRAS/355/20 VIII/79 Kassin S.A.: J/ApJS/162/80 Kassis M.: J/AJ/113/1723 J/AJ/130/1763 Kastner J.: J/ApJS/160/319 Kastner J.H.: J/A+A/275/163 J/ApJ/549/441 J/ApJS/160/511 Kasun S.: J/A+A/459/679 Katajainen S.: J/A+A/427/769 J/A+AS/129/577 J/A+AS/130/109 J/A+AS/143/357 Katgert J.K.: J/A+A/35/393 Katgert P.: J/A+A/310/31 J/A+A/310/8 J/A+A/446/19 J/A+AS/129/399 J/A+AS/58/1 KATGERT P.: J/A+AS/93/399 Kato D.: J/ApJ/596/1064 J/MNRAS/353/705 J/MNRAS/354/980 J/MNRAS/368/534 J/MNRAS/370/1979 Kato K.: J/A+A/420/673 Kato M.: J/ApJS/113/121 Kato T.: II/273 J/A+A/460/331 Kato Y.: J/MNRAS/330/137 Katsanis R.: J/AJ/112/1335 Katsiyannis A.C.: J/MNRAS/342/1241 Katsova M.M.: J/A+AS/139/555 Katz D.: J/A+A/404/661 J/A+A/412/495 J/A+A/426/1007 J/A+AS/133/221 Kaufer A.: II/202 KAUFER A.: J/A+A/289/740

Kaufer A.: J/A+A/389/931 J/A+A/437/467 J/A+A/457/265 J/A+A/459/423 J/A+AS/133/299

J/AJ/127/2723 Kauffmann G.: J/A+A/444/365 J/ApJS/157/175 J/MNRAS/362/9 Kaufl H.-U.: J/A+A/473/369 Kaufman M.: J/AJ/121/182 Kaufman V.: VI/74 Kaufman V.; Zalubas R.: VI/73 Kausch W.: J/A+A/413/1037 Kautsch S.J.: J/A+A/445/765 Kavars D.W.: J/ApJ/626/887 Kawadu T.: J/MNRAS/370/1979 Kawaguchi N.: J/PASJ/52/631 J/PASJ/52/997 Kawaguchi S.: J/PASJ/52/631 Kawai E.: J/PASJ/52/631 J/PASJ/52/997 Kawai H.: J/other/APh/26.282 Kawai N.: J/other/GCN/1070.1 Kawakita H.: J/PASJ/56/295 Kawamura A.: J/PASJ/56/313 Kawanomoto S.: J/A+A/468/663 J/PASJ/57/13 J/PASJ/57/45 J/PASJ/59/245 Kawara K.: J/A+A/365/409 J/A+A/413/843 KAWARA K.: VII/60 Kawasaki W.: J/AJ/123/1200 Kawka A.: J/MNRAS/335/673 Kaye A.B.: J/ApJ/525/420 Kayo I.: J/MNRAS/382/412 Kayser R.: J/A+A/318/680 Kazarian M.A.: J/other/Ap/44.335 J/other/Ap/48.344 Kazarovets E.V.: II/172B II/195B II/205 II/219 J/other/IBVS/5721 J/other/PZ/23.141 J/PAZh/28/201 KAZES I.: J/A+A/299/347 Kazes I.: J/A+A/411/381 J/A+AS/104/529 J/A+AS/116/193 J/A+AS/80/215 J/MNRAS/307/236 Kazlanskas A.: II/93 KAZLAUSKAS A.: II/157A II/196 Kazlauskas A.: J/BaltA/11/219 J/BaltA/11/231 J/BaltA/12/323 J/BaltA/13/1 J/BaltA/13/555 J/BaltA/15/483Short

J/BaltA/7/369 Keane M.J.: II/162 Kedziora-Chudczer L.: J/MNRAS/369/449 Kedziora-Chudczer L.L.: J/MNRAS/325/1411 Keeble O.: VII/221 Keefe H.: V/68A KEEL W.C.: J/A+AS/108/61

Keel W.C.: J/A+AS/96/389 J/AJ/109/14 J/AJ/111/696 J/AJ/129/26 J/AJ/129/31 J/ApJ/581/844

J/ApJ/613/841 J/ApJS/106/27 J/ApJS/146/267 J/ApJS/158/139 VI/76 Keenan F.P.: J/A+A/347/69 J/A+A/367/86 J/A+A/399/799 J/A+A/407/769 J/A+A/410/349 J/A+A/416/383

J/A+A/418/371 J/A+A/419/1123 J/A+A/420/783 J/A+A/424/363 J/A+A/427/763 J/A+A/429/1117 J/A+A/432/1151 J/A+A/433/745 J/A+A/436/1141 J/A+A/439/1215 J/A+A/441/831 J/A+A/450/1249 J/A+A/452/503 J/A+A/455/1157 J/A+A/460/331 J/A+A/460/959 J/A+A/463/399 J/A+A/473/995 J/A+A/475/393 J/A+AS/139/135 J/ApJS/123/311 J/MNRAS/367/1478 J/MNRAS/372/1117 J/MNRAS/379/647 J/MNRAS/379/816

Keenan P.C.: III/150A III/66 J/ApJ/518/859 J/ApJS/105/419 J/ApJS/43/379 Kegel W.H.: J/A+AS/114/175 J/A+AS/142/113 Kehoe R.: J/AJ/119/1901 J/AJ/127/2436 Kehrig C.: J/A+A/423/133 Keinanen P.: J/A+A/427/769 J/A+AS/132/305 Keipert S.: J/A+A/466/255 Keller S.C.: J/A+AS/134/489 J/AJ/119/1748 J/AJ/122/248 J/AJ/124/2039

Kellermann K.I.: J/A+A/342/378 J/A+A/386/97 J/A+A/399/469 J/AJ/108/1163 J/AJ/123/2402

J/AJ/130/2473 J/AJ/77/797 J/ApJ/609/539 J/MNRAS/358/1159 Kellett B.J.: J/MNRAS/274/1194 Kelley J.L.: J/other/APh/26.282 Kelley R.: J/ApJS/96/303 Kellogg E.: IX/3 Kelly D.: J/ApJS/154/155 Kelly D.M.: J/AJ/131/250 Kelm B.: J/A+A/335/912 J/A+A/391/35 J/A+A/456/467 J/AJ/129/1832 J/MNRAS/364/146 Kels A.P.: VIII/82 Kelson D.: J/ApJ/523/540 Kelson D.D.: J/ApJ/463/26 J/ApJ/500/714 J/ApJ/508/491 J/ApJ/512/48 J/ApJ/514/614 J/ApJ/521/155

J/ApJ/529/723 J/ApJ/661/750 Kembhavi A.: J/A+A/465/1099 J/AJ/118/1435 Kemp J.: J/A+A/466/1031 Kendall T.R.: J/A+A/448/189 J/MNRAS/374/372 Kenderdine S.: VIII/30 Kendig D.J.: J/ApJS/101/259 J/ApJS/107/227 Kendziorra E.: J/A+A/443/103 Kennea J.: J/A+A/448/L9 J/MNRAS/369/156 Kennedy P.M.: III/78 Kennefick: J/ApJS/119/189 Kennefick J.D.: J/AJ/114/2269 Kenney J.D.P.: J/AJ/123/1411 Kennicutt Jr.R.C.: J/ApJ/463/26 Kennicutt R. C.: J/ApJ/510/104 Kennicutt R.C.: J/ApJ/514/614 J/ApJ/521/155 J/ApJ/523/540 Kennicutt R.C. Jr.: J/ApJS/73/661 Kennicutt R.C.JR: J/AJ/132/231 Kennicutt R.C.Jr: J/ApJ/507/655 J/ApJ/508/491 J/ApJ/512/48 J/ApJ/529/723 J/ApJ/633/871 Kennicutt R.C.JR: J/ApJS/164/81 J/ApJS/165/307 Kennicutt R.C.Jr: VII/141 Kennicutt R.JR: J/AJ/130/1324 Kent B.R.: J/AJ/130/2613 VIII/77

Kent S.: J/AJ/121/2308 J/AJ/121/2331 J/AJ/123/567 J/AJ/126/1023 J/AJ/126/2579 J/ApJ/588/824

J/ApJS/163/80 Kent S.M.: J/AJ/105/393 J/AJ/130/873 Kenter A.: J/ApJS/161/9 Kenter A.T.: J/ApJ/641/140 Kenworthy M.A.: J/ApJS/161/154 Kenyon M.J.: J/MNRAS/356/89 Kenyon S.J.: J/AJ/121/2673 J/AJ/121/974 J/AJ/122/1289 J/AJ/126/1362 J/AJ/130/1097 J/ApJ/582/668

J/ApJ/636/1002 J/ApJS/101/117 J/ApJS/112/109 KEPNER J.: J/ApJ/438/813 Kerber F.: J/A+A/400/877 J/A+A/408/1029 J/A+A/470/317 J/A+AS/117/369 Kerins E.: J/MNRAS/351/1071 Kern K.: J/AJ/129/393 Kerr T.H.: J/MNRAS/373/781 KERSCHBAUM F.: II/177 Kerschbaum F.: J/A+A/338/581 J/A+A/375/527 J/A+A/403/93 J/A+A/406/51 J/A+A/427/613 KERSCHBAUM F.: J/A+AS/106/397 Kerschbaum F.: J/A+AS/113/441 J/A+AS/130/455 KERSCHBAUM F.: J/A+AS/99/31 Kerton C.R.: J/A+A/399/1083 J/AJ/122/3155 J/AJ/131/1501 J/ApJS/126/85 J/ApJS/144/47 J/MNRAS/373/1203

J/MNRAS/379/289 Kervella: J/A+A/428/587 Kervella P.: J/A+A/453/155 J/A+A/453/309 J/A+A/459/669 J/A+A/464/373 J/A+A/476/73 KESKIN V.: J/A+A/281/811 J/A+A/287/817 Keskin V.: J/A+A/336/920 Kessler M.F.: J/A+A/359/865 J/A+A/365/545 J/A+A/407/791 Kestel M.: J/A+A/390/39 J/A+A/406/L9 J/A+A/410/813 J/other/APh/26.282 Kester D.: J/A+A/390/1033 Kesteven M.J.: J/AJ/124/1954 J/AJ/128/16 J/MNRAS/289/225 J/MNRAS/301/1019 J/MNRAS/354/305

J/MNRAS/371/898 Ketsaris N.A.: J/AZh/80/535 Kettenis M.M.: J/A+A/472/805 Kettle D.: J/A+A/474/1093 Kettler J.: J/A+A/390/39 J/A+A/410/813 Kewley L.: J/ApJ/595/685 J/ApJ/633/871 J/ApJS/139/369 J/ApJS/155/271 Kewley L.J.: J/AJ/127/2002 J/ApJ/617/240 J/ApJ/619/755 J/ApJS/132/37 J/PASP/117/227 Keyes C.D.: J/ApJ/636/1002 Khabrakhmanov A.: J/AZh/72/12 Khalisi E.: I/264 I/286 Khanzadyan T.: J/A+A/426/171 J/A+A/447/609 Kharchenko N.: J/A+A/384/925 J/A+AS/137/305 J/AN/316/225 Kharchenko N.V.: I/280A III/239 III/254 J/A+A/384/145 J/A+A/438/1163 J/A+A/440/403 J/A+A/468/151

J/A+A/477/165 J/A+AS/134/525 J/AN/325/740 J/AN/328/889 J/AZh/80/291 J/BaltA/9/646 V/126 Kharitonov A.V.: III/186 III/202 J/A+AS/92/1 J/AZh/73/67 J/AZh/73/906 J/AZh/75/197 J/AZh/75/70

J/AZh/77/619 KHARRAZI M.: J/A+A/293/75 Kholikov S.: J/A+A/390/717 KHOLOPOV P.N. (ed.): II/137 Kholopov P.N. Efremov Y.N.: II/69 Kholopov P.N.: II/12 II/13 II/139B II/140 II/172B II/195B II/205 II/214A II/70 KHOMYAKOVA M.B.: I/18 Khovritchev M.Yu.: I/290 Khristoforova M.: J/A+A/431/773 KHRUTSKAYA E.V.: I/137 Khrutskaya E.V.: I/290 I/309 Khruzina T.S.: III/208 Kiasatpoor A.: V/31A Kidd T.T.: J/AJ/128/712 Kidger M.: J/A+AS/125/399 Kidger M.R.: J/AJ/122/2055 J/AJ/125/3311 Kiehling R.: III/124 Kienlin A. von: J/A+A/438/1175 Kienzle F.: J/A+A/337/779 J/A+A/341/818 Kiger J.R.: J/AJ/121/97 Kijak J.: J/A+AS/147/195 Kikuchi S.: J/A+A/327/145 Kilborn V.A.: J/AJ/123/873 J/AJ/124/1954 J/AJ/124/690 J/AJ/128/16 J/ApJS/165/307 J/MNRAS/337/641

J/MNRAS/342/738 J/MNRAS/369/1351 J/MNRAS/370/1223 Kilgard R.E.: J/MNRAS/357/401 Kilic M.: J/AJ/128/1825 J/AJ/131/582 Kilkenny D.: III/137 J/AJ/123/2822 J/AJ/126/3007 J/AJ/128/463 J/MNRAS/236/207 J/MNRAS/287/867

J/MNRAS/320/451 J/MNRAS/325/931 J/MNRAS/334/20 J/other/SAAOC/14.11 J/other/SAAOC/15.40 J/other/SAAOC/15.53 J/other/SAAOC/15.65 J/other/SAAOC/15.78

Killeen N.E.B.: J/A+A/366/481 J/A+AS/122/79 J/A+AS/124/509 J/MNRAS/352/1245 Kilmartin P.M.: J/MNRAS/330/137 J/MNRAS/331/609 Kilpio E.: J/A+A/384/925 Kilpio E.Yu.: J/BaltA/9/646 Kim A.: J/A+A/389/149 J/A+A/421/509 Kim A.G.: J/A+A/437/789 Kim B.G.: J/PASJ/56/313 Kim C.: J/A+A/405/231 J/A+A/430/255 KIM C.: J/AJ/106/2493 Kim C.: J/MNRAS/359/865 Kim C.-H.: J/AJ/128/846 J/MNRAS/379/1665 Kim D.-C.: J/AJ/126/1607 J/ApJ/522/113 J/ApJS/143/277 J/ApJS/143/315 J/PASP/110/779 Kim D.-W.: J/AJ/131/527 J/ApJ/442/523 J/ApJ/586/826 J/ApJ/618/123 J/ApJ/644/829 J/ApJ/645/955

J/ApJS/150/19 J/ApJS/161/271 KIM D.-W.: J/ApJS/70/329 Kim D.-W.: J/ApJS/80/531 J/ApJS/97/141 Kim H.-G.: J/ApJS/165/360 KIM H.-I.: J/A+A/291/110 Kim H.-I.: J/A+A/437/61 Kim H.I.: J/A+A/445/69 Kim H.-I.: J/A+A/454/717 J/AJ/126/2840 Kim H.-R.: J/ApJS/165/360 Kim J.: J/PASP/118/636 Kim J.H.: J/ApJ/632/L131 Kim J.S.: J/PASP/118/1690 Kim J.-W.: J/A+A/454/717 Kim K.-M.: J/PASP/118/636 Kim K.-S.: J/A+A/404/621 KIM K.-T.: J/A+AS/105/385 Kim K.-T.: J/A+AS/105/403 Kim M.: J/ApJ/644/829 J/ApJ/645/955 Kim R.S.J.: J/AJ/123/1807 J/ApJS/148/243 Kim S.C.: J/AJ/119/777 J/AJ/126/2840 Kim S.-J.: J/ApJS/165/360 Kim S.-L.: J/A+A/404/621 J/A+A/405/231 J/A+A/459/557 Kim S.S.: J/ApJ/581/258 Kim T.-S.: J/A+A/373/757 J/ApJ/596/768 J/MNRAS/335/555 Kim Y.-C.: J/A+A/454/717 Kim Y-C.: J/AJ/112/2408 Kimble R.A.: J/AJ/127/2738 J/AJ/128/1990 J/AJ/130/1693 J/ApJ/532/308 J/ApJ/612/202 J/ApJS/150/1 Kimeridze G.N.: J/A+A/402/151 Kimeswenger S.: J/A+A/363/62 J/A+A/387/1 J/A+A/413/1037 J/A+AS/135/133 J/other/RMxAA/37.115 Kimura M.: J/MNRAS/357/1357 J/PASJ/55/1079 Kindl E.: VII/213 Kinemuchi K.: J/AJ/127/2436 King A.R.: J/ApJ/609/735 KING C.R.: VI/40 King D.S.: I/86A King E.A.: J/AJ/110/880 J/PASJ/55/351 King I.R.: J/A+A/287/769 J/A+A/288/413 J/A+A/391/945 J/AJ/109/218 J/AJ/109/639 King J.R.: J/ApJ/602/L117 J/ApJ/632/L131 King N.L.: J/AJ/127/1632 Kingsburgh R.L.: J/AJ/116/2246 Kinman T.: J/A+A/364/102 J/A+A/391/1039 Kinman T.D.: J/AJ/108/1722 J/MNRAS/375/1381 KINNEY A. L.: III/167 Kinney A.L.: III/157 J/AJ/110/2739 Kinney E.K.: J/AJ/124/1651 Kinnunen T.: IV/23 Kinoshita H.: J/A+A/312/1017 Kippen R. M.: J/ApJS/120/399 Kippen R.M.: IX/20A J/A+AS/143/145 Kipper M.: J/A+A/276/389 Kipper T.: J/A+A/276/389 Kireeva N.N.: II/139B II/140 II/172B II/205 J/other/IBVS/5721 Kirhakos S.: J/ApJ/409/28 J/PASP/106/646 Kirkman D.: J/AJ/119/2571 Kirkpatrick H.: J/ApJS/105/93

Kirkpatrick J.D.: II/246 J/AJ/125/3302 J/AJ/126/2421 J/AJ/126/2487 J/AJ/127/2856 J/AJ/128/463

J/AJ/133/439 J/ApJ/476/311 J/ApJ/564/421 J/ApJ/637/1067 J/ApJS/94/749 J/other/ARA+A/43.195 Kirshner R.P.: J/A+A/428/555 J/AJ/130/2453 J/AJ/131/1648 J/AJ/131/527 J/AJ/83/1549 J/AJ/88/1285

J/AJ/95/1629 J/ApJ/314/493 J/ApJ/594/1 J/ApJ/602/571 J/ApJ/607/665 J/ApJ/626/L5 J/ApJ/627/579 J/ApJS/120/247 J/ApJS/130/237 J/ApJS/99/223 VII/203

Kiselev A.A.: I/292 J/PAZh/27/456 Kiseleva T.P.: J/A+AS/139/47

Kisielius R.: J/A+A/387/1135 J/A+A/416/383 J/A+A/420/783 J/A+A/424/363 J/A+A/433/745

J/A+A/455/1157 J/A+AS/133/257 J/A+AS/135/359 J/A+AS/137/157 J/A+AS/142/85 Kislyuk V.: I/261 Kiss C.: J/A+A/410/879 J/A+A/418/131 Kiss Cs.: J/A+A/463/1227 Kiss L.L.: J/A+A/370/496 J/A+A/374/712 J/A+A/375/285 J/A+A/376/561 J/A+A/390/585 J/A+A/394/943

J/A+A/394/97 J/A+A/410/879 J/A+A/461/613 J/A+A/463/589 J/A+A/465/943 J/other/P+SS/53.925 Kiss Z.T.: J/A+A/453/923 J/A+A/463/1227

Kissler-Patig M.: J/A+A/415/123 J/A+A/435/65 J/A+A/459/391 J/A+A/464/495 J/A+A/472/111

J/A+AS/134/59 J/A+AS/134/75 J/MNRAS/343/665 J/MNRAS/369/697 Kitamoto S.: IX/19 Kitamura N.: J/other/APh/26.282 Kitsionas S.: J/A+A/434/475 J/A+A/459/693 J/MNRAS/352/91 Kiyaeva O.: J/A+A/448/1075 Kiyaeva O.V.: I/292 J/PAZh/24/868 J/PAZh/27/456 Kjaergaard: J/MNRAS/273/1097 KJAERGAARD P.: III/79 Kjaergaard P.: J/A+A/470/39 J/ApJ/542/673 Kjeldsen H.: J/A+A/439/387 J/A+A/459/557 J/A+A/465/965 J/A+AS/140/135 Kjfrgaard P.: J/A+A/387/26 Kjurkchieva D.P.: J/A+A/424/993 Klaas U.: J/A+A/359/865 J/A+A/401/873 J/A+A/422/39 J/A+A/466/1205 Klagyivik P.: J/A+A/426/1001 J/A+A/445/617 Klamer I.: J/MNRAS/347/837 Klare G.: II/34A II/62 Klaric M.: J/AJ/121/182 Klassen A.: J/A+AS/119/489 Klein R.: J/ApJS/161/361 Klein S.R.: J/other/APh/26.282 Klein U.: J/A+A/406/579 J/A+A/435/863 J/A+A/451/431 J/A+AS/111/311 J/A+AS/114/21 J/A+AS/120/77

J/A+AS/121/321 J/A+AS/127/119 J/A+AS/133/129 J/A+AS/139/359 Kleinheinrich M.: II/253 J/A+A/377/442 Kleinman S.J.: J/A+A/444/643 J/AJ/123/3409 J/AJ/124/1651 J/AJ/126/1023 J/AJ/131/1674 J/AJ/132/676

J/ApJ/607/426 J/ApJS/165/1 J/PASJ/55/771 Kleinmann S.G.: II/126 III/114 KLEINMANN S.G.: J/ApJ/411/188 Klemola A.R.: I/11 KLEMOLA A.R.: I/140 Klemola A.R.: I/152 I/153 I/199A I/200 I/214 I/283A I/293 I/76 KLEMOLA A.R.: I/77 Klemola A.R.: J/A+AS/124/157 J/AJ/124/601 Klepser S.: J/other/APh/26.282 Kleyna J.: J/ApJ/635/290 J/MNRAS/330/792 Kleyna J.T.: J/AJ/131/527 J/AJ/131/895 Klimek E.S.: J/A+A/371/79 KLINGLESMITH D.A.: VII/94

Klochkova V.G.: III/243 III/245 J/A+A/381/32 J/A+A/397/1035 J/A+A/425/671 J/A+AS/131/469 J/A+AS/147/5 J/AZh/81/1104 J/AZh/81/333 J/AZh/83/265 J/PAZh/25/30 J/PAZh/26/115 J/PAZh/26/465 J/PAZh/26/510

Klochkova V.G..: J/PAZh/27/381 Klochkova V.G.: J/PAZh/27/841 J/PAZh/32/737 Klock B.: V/31A Klockner H.-R.: J/MNRAS/373/1531 Klotz A.: J/A+A/425/927 J/AJ/133/1470 J/other/IBVS/5767 KLUKE J.: VII/5A Klutsch A.: J/A+A/454/301 Klypin A.: J/ApJ/598/260 Klypin A.A.: J/ApJ/645/228 Knapen J.H.: J/A+A/414/23 J/A+A/459/L13 J/A+A/466/905 KNAPEN J.H.: J/AJ/106/56 Knapen J.H.: J/MNRAS/297/255 J/MNRAS/317/234 VI/112 Knapik A.: J/A+A/321/236 J/A+A/321/L9 J/A+A/332/L53 J/A+A/342/773 J/A+A/344/263 J/A+A/369/178

J/A+A/390/967 Knapp G.: J/A+A/399/1167

Knapp G.R.: J/A+A/403/993 J/A+A/423/755 J/A+A/444/643 J/A+AS/129/363 J/AJ/121/2308 J/AJ/121/31 J/AJ/122/503 J/AJ/123/3409 J/AJ/123/567 J/AJ/124/1651 J/AJ/125/1711 J/AJ/125/1817 J/AJ/125/3302 J/AJ/126/1023 J/AJ/126/2579 J/AJ/127/2838 J/AJ/127/3553 J/AJ/127/899 J/ApJ/607/426 J/ApJS/165/1

KNAPP G.R.: J/ApJS/70/329 Knauth D.C.: J/A+A/408/305 Knebe A.: J/A+A/459/L1 J/A+A/462/811 Knee L.B.G.: J/A+AS/116/157 J/AJ/122/3155 J/AJ/131/1501 KNEER R.: II/200 Kneer R.: J/ApJS/117/319 Kneib J.-P.: B/bax J/A+A/372/391 J/A+A/407/791 J/ApJ/591/53 J/MNRAS/330/1 Kneissl R.: J/MNRAS/341/1066 J/MNRAS/360/340 Knerr E.J.: J/AJ/127/899 Knezek P.M.: J/AJ/123/873 J/AJ/124/1954 J/AJ/124/690 J/AJ/128/16 J/ApJS/165/307 J/MNRAS/337/641

J/MNRAS/346/787 J/MNRAS/355/1303 Kniazev A.: J/A+A/419/469 J/AJ/121/66 J/AJ/123/1292 J/AJ/130/968

Kniazev A.Y.: J/A+A/366/771 J/A+A/373/24 J/A+A/374/907 J/A+A/442/109 J/A+AS/142/417

J/AJ/127/704 J/ApJS/153/429 Kniazev A.Yu.: J/A+AS/135/511 J/A+AS/137/299 Kniazeva T.F.: J/A+A/366/771 J/A+A/374/907 J/AJ/127/704 Kniffen D.A.: IX/9 J/ApJS/101/259 J/ApJS/107/227 J/ApJS/123/79 J/ApJS/94/551 Knigge C.: J/ApJ/625/156 Knoedlseder J.: J/A+AS/143/145 Knop R.A.: J/A+A/437/789 Knopp G.P.: J/ApJS/109/367 KNOWLES S.H.: VII/50 Knox E.R.: J/AJ/132/161 Knox Jr J.R.: J/A+AS/103/1 Knude J.: IV/25 J/A+A/372/276 J/A+A/463/1029 J/A+AS/130/477 J/A+AS/92/841 Knyazeva L.N.: III/186 III/202 J/A+AS/92/1 J/AZh/73/67 J/AZh/73/906 J/AZh/75/197 J/AZh/75/70

J/AZh/77/619 KO P.W.: J/AJ/104/1706 Kobayashi H.: J/PASJ/52/631 J/PASJ/52/997 Kobayashi: J/ApJ/527/573 Kobayashi N.: J/AJ/125/1537 J/AJ/130/2212 J/ApJ/566/974 J/ApJS/118/177 J/PASJ/57/165 Kobayashi Y.: J/ApJS/141/23 J/ApJS/141/31 Kober G.V.: J/ApJS/157/138 KOBULNICKY H.A.: J/AJ/107/1433 Kobulnicky H.A.: J/AJ/110/268 J/ApJ/559/864 J/ApJ/599/1006 J/ApJ/617/240 J/ApJ/635/560 Kocer D.: J/A+A/371/1078 J/A+A/400/1043 J/A+A/406/975 J/MNRAS/316/514 Koch A.: J/A+A/381/500 J/AJ/131/895 Koch D.: IX/3 Koch Miramond L.: J/A+A/384/504 Koch R.H.: J/AJ/120/1548 J/MNRAS/379/1665 KOCH R.H.: VI/23 Koch R.H.: VI/44 KOCH S.J.R.H.: J/A+AS/102/343 Kochanek C.S.: J/AJ/130/1977 J/AJ/131/527 J/ApJ/560/566 J/ApJ/625/633 J/ApJ/641/140 J/ApJS/161/9 Kochiashvili N.T.: J/AZh/78/408 Kochukhov O.: J/A+A/374/265 J/A+A/450/763 Kodaira K.: J/ApJS/118/177 J/PASJ/56/1025 J/PASJ/57/165 VII/147 Kodama T.: J/ApJ/528/123 J/ApJ/634/861 J/MNRAS/323/839 J/MNRAS/357/1357 J/MNRAS/365/1392

J/MNRAS/379/1546 J/PASJ/55/1079 J/PASJ/57/165 Koechlin L.: J/A+A/367/865 J/A+AS/134/545 J/ApJS/131/561 J/MNRAS/367/1170 Koehler R.: J/A+A/356/541 J/A+A/369/249 J/A+A/376/982 J/A+A/458/461 Koehler S.: J/A+AS/128/507 Koehler T.: J/A+A/358/77 J/A+AS/115/235 Koekemoer A.: J/AJ/131/1216 J/ApJ/613/200 J/ApJS/139/369 Koekemoer A.M.: J/AJ/128/1974 J/ApJ/595/685 J/ApJS/161/21 Koelbloed D.: J/other/BAN/14.265 Koen C.: J/MNRAS/287/867 J/MNRAS/320/451 J/MNRAS/331/45 J/MNRAS/334/20 Koenigsberger G.: J/A+A/347/583 Koenoenen P.: J/A+A/383/197 Koepke L.: J/other/APh/26.282 KOEPPEN J.: III/131 Koeppen J.: J/A+A/282/436 J/A+A/307/215 J/A+A/463/265 Koerner D.W.: J/AJ/125/3302 Koester D.: J/A+A/363/1040 J/A+A/378/556 J/AJ/132/676 J/ApJ/607/426 Koff R.: J/A+A/465/331 Kogoshvili N.: J/A+AS/91/371 KOGOSHVILI N.G.: VII/40 VII/90A KOGURE T.: III/177 III/177 III/177 III/177 Kogut A.: J/ApJS/148/97 Kohl-Moreira J.L.: J/AJ/130/425 Kohnen G.: J/other/APh/26.282 Kohnle A.: J/A+A/390/39 J/A+A/406/L9 J/A+A/410/813 Kohno M.: J/ApJ/567/423 Kohoutek L.: III/205 IV/24 J/AN/318/35 J/AN/323/484 J/AN/324/437 Kohyama Y.: J/ApJS/102/411 J/ApJS/63/661 Koike C.: J/A+A/449/583 Koivisto P.: J/A+A/427/769 J/A+A/440/409 Kojo K.: J/ApJS/74/291 Kojoian G.: J/A+AS/147/169 J/ApJS/98/369 Kokko M.: J/A+A/350/513 J/A+A/362/223 Kokkonen K.: J/A+AS/132/305 Kokkotas K.D.: J/A+A/366/565

Kolaczkowski Z.: J/A+A/345/505 J/A+A/376/144 J/A+A/398/541 J/A+A/478/659 J/AcA/53/27

J/MNRAS/298/753 Kolanoski H.: J/other/APh/26.282 Kolatt T.: VII/198 Kolb E.W.: J/ApJ/596/437 Kolb U.: B/cb V/110 V/113D V/123A KOLB U.: V/82 Kolb U.: V/99 Kolenberg K.: J/A+A/463/243 KOLESNIK L.N.: II/132 Kolesnik L.N.: III/230 Kolesnik Y.B.: I/241 KOLESNIK Y.B.: J/A+AS/103/427 Kolev D.: J/A+AS/116/1 J/A+AS/122/43 Koleva M.: III/251 Kolkman O.M.: J/A+AS/119/37 Kollatschny W.: J/A+A/314/43 J/A+A/475/487 J/ApJ/510/659 Kollgaard R.: J/A+AS/134/221 Kollgaard R.I.: J/A+A/323/739 J/A+AS/122/235 J/ApJ/525/127 KOLLGAARD R.I.: J/ApJS/93/145 Kollmeier J.A.: J/AJ/130/1916 J/ApJS/152/103 Kolokotronis V.: J/MNRAS/349/135 Kolotilov E.A.: III/208 J/A+A/311/484 KOLOTILOV E.A.: J/A+AS/112/457 Kolotilov E.A.: J/AZh/71/618 Kolpak M.A.: J/ApJ/582/756

Komissarova G.V.: J/PAZh/27/841 J/PAZh/29/858 J/PAZh/30/855 J/PAZh/32/48 J/PAZh/32/662

J/PAZh/32/737 J/PAZh/33/679 Komiyama Y.: J/AJ/127/180 J/ApJS/137/279 J/MNRAS/357/1357 J/PASJ/55/1079 J/PASJ/56/1025

J/PASJ/57/165 Kommers J.: J/ApJS/156/217 Kommers J.M.: J/ApJS/134/385 Komossa S.: J/A+A/310/384 J/A+A/456/953 J/A+A/459/55 J/ApJ/590/73 J/ApJS/129/435 Kompaneets D.: J/ApJ/563/80 Konacki M.: J/AJ/130/2866 J/ApJ/636/1020 Kondo M.: II/106 II/106 Kondo T.: J/PASJ/52/631 Kondrat’eva L.N.: VI/93 Kondratiev V.B.: J/A+A/350/513 J/A+AS/139/513 Kondratyeva L.N.: III/203 Kong A.K.H.: J/ApJ/577/738 J/ApJ/599/1067 J/ApJ/609/735 J/ApJ/610/247 J/ApJ/636/1002 Kong M.-Z.: J/ApJ/645/890 Kong X.: J/A+A/389/845 J/A+A/396/503 J/AJ/126/1286 Kong Xu: J/A+A/425/417 Konichek V.V.: J/A+A/420/447 Konin V.V.: I/309 Konnikova: J/AZh/72/291 Konnikova V.K.: J/AZh/77/407 J/AZh/80/978 J/AZh/82/420 J/AZh/83/241 VIII/49 Kononen P.: J/A+A/427/769 J/A+AS/132/305 Konopelko A.: J/A+A/390/39 J/A+A/406/L9 J/A+A/410/813 Kontinen S.: J/A+A/383/197

Kontizas E.: J/A+A/369/932 J/A+A/405/111 J/A+A/447/89 J/A+A/451/431 J/A+AS/136/525

J/MNRAS/351/1290 VII/183 Kontizas M.: J/A+A/369/932 J/A+A/405/111 J/A+A/423/919 J/A+A/447/89 J/A+A/451/431

J/MNRAS/351/1290 VII/183 Konyves V.: J/A+A/417/745 J/A+A/463/1227 Koo D. C.: J/ApJS/122/355

Koo D.C.: J/A+AS/58/39 J/AJ/127/3121 J/AJ/90/418 J/ApJ/481/673 J/ApJ/505/50 J/ApJ/571/136 J/ApJ/599/1006 J/ApJ/625/6 J/ApJS/109/45 J/ApJS/142/1 J/ApJS/157/175 J/ApJS/159/41 J/MNRAS/317/801 VII/200

Koopmann R.A.: J/AJ/112/2408 J/AJ/130/2613 J/AJ/131/716 Koopmans L.V.E.: VIII/72 Kopacki G.: J/A+A/358/547 J/A+A/369/862 J/A+A/376/144 J/A+A/398/541 J/AcA/53/27 Kopan E.L.: II/246 J/AJ/128/787 KOPAN E.L.: VII/109 Kopan G.: II/156A Kopff A.: I/143 I/15 KOPFF A.: I/24 Kopon D.A.: J/AJ/127/501 Kopp M.: J/A+AS/122/415 Koppen J.: V/127 Kopylov A.I.: J/A+A/382/389 J/A+A/403/555 J/A+A/421/455 J/AZh/72/613 J/AZh/73/835 J/AZh/79/589 KOPYLOV A.I.: J/MNRAS/243/390 Kopylov A.I.: J/other/BSAO/40.5 J/PAZh/27/170 J/PAZh/27/403 Kopylova F.G.: J/A+A/382/389 J/A+A/403/555 J/A+A/421/455 J/PAZh/27/170 J/PAZh/27/403 Korakitis R.: J/A+A/405/111 Koranyi D.M.: J/AJ/116/2108 J/AJ/119/44 J/AJ/123/100 J/AJ/131/527 Koratkar A.: J/ApJS/143/257 Koratkar A.P.: J/ApJS/105/93 Korcakova D.: J/A+A/419/607 J/A+A/459/849 Korhonen H.: J/A+A/374/1049 J/A+A/390/179 J/A+A/476/881 Koribalski B.: J/AJ/123/873 J/AJ/124/690 J/AJ/129/220 J/ApJS/165/307 J/MNRAS/337/641 Koribalski B.S.: J/AJ/124/1954 J/AJ/128/16 Kormendy J.: J/AJ/129/2138 Korn A.J.: J/A+A/389/931 J/A+A/437/467 Kornilov E.V.: I/242 I/266 Kornilov V.G.: J/A+AS/139/47 Kornmeyer H.: J/A+A/390/39 J/A+A/406/L9 KOROBOVA V.A.: I/18 Korotin S.A.: J/A+A/418/551 J/AZh/80/458 Korotkova G.G.: J/A+A/433/751 Korpi M.: J/A+AS/132/305 Kortazzi I.: I/310 Korzennik S.G.: J/A+A/410/1051 Koshut T.M.: IX/20A Koski K.: J/AJ/119/2801 Koskimies M.: J/A+A/427/769 Kostjuk N.D.: IV/27 Kotak R.: J/MNRAS/369/1880 Kotanyi C.G.: J/A+AS/41/421 Kothes R.: J/A+A/363/141 J/A+AS/131/11 Kotilainen J.: J/A+AS/116/157 KOTILAINEN J.: J/ApJ/413/453 Kotilainen J.K.: J/AJ/127/3075 J/MNRAS/351/1290 Kotiranta M.: J/AJ/133/1947 Kotnik-Karuza D.: J/A+A/381/507 J/A+A/452/503 Kotoneva E.: J/A+AS/143/357 J/MNRAS/336/879 Kouba Y.: J/MNRAS/367/1609

Koubsky P.: J/A+A/304/235 J/A+A/309/521 J/A+A/319/867 J/A+A/324/965 J/A+A/328/551 J/A+A/419/607 J/A+A/446/583 J/A+A/455/1037 J/A+A/459/849 J/A+A/463/1061 J/A+A/463/233

KOUBSKY P.: J/A+AS/107/403 Koubsky P.: J/A+AS/125/75 Koulis C.: J/A+AS/121/223 Koulouridis E.: J/ApJ/639/37 J/MNRAS/360/782 Kouveliotou C.: IX/20A J/ApJ/563/L49 J/ApJ/615/887 J/ApJ/634/L89 J/ApJS/120/399 J/ApJS/134/385

J/ApJS/156/217 Kouwenhoven M.B.N.: J/A+A/430/137 J/A+A/464/581 Kovachev B.: J/A+A/376/745 J/A+AS/117/261 Kovacs G.: J/A+A/312/111 J/A+A/363/L1 J/A+A/371/579 J/A+A/454/257 J/ApJ/542/257 J/ApJ/598/597 Kovalenko A.V.: J/AZh/71/110 Kovalev Y.A.: J/PASJ/52/997 Kovalev Yu.A.: J/A+AS/139/545 Kovalev Y.Y.: J/A+AS/139/545 J/AJ/129/1163 J/AJ/130/2473 J/AJ/131/1872 J/ApJ/609/539 Kovaleva D.A.: J/AZh/78/1104 J/AZh/79/259 Kovalevsky J.: J/A+A/331/81 KOVO O.: J/A+A/299/89 Kovtun: IX/36 Kovtyukh V.: J/MNRAS/359/865 Kovtyukh V.V.: J/A+A/281/465 Kovtyukh V.V: J/A+A/305/551

Kovtyukh V.V.: J/A+A/381/32 J/A+A/392/491 J/A+A/411/559 J/A+A/418/551 J/A+A/427/933

J/AJ/129/433 J/AJ/130/1880 J/AZh/80/458 J/AZh/84/429 J/AZh/86/158 J/MNRAS/371/879 KOWAL C.T.: II/103 Kowal C.T.: VII/4A KOWALSKI M.P.: J/ApJS/56/403 Kox H.: J/AN/283/109 Koyama K.: J/AJ/125/1537 J/ApJ/566/974 J/ApJ/567/423 J/ApJS/134/77 J/PASJ/55/161 J/PASJ/55/635

J/PASJ/55/653 J/PASJ/57/135 Koyama Y.: J/AJ/130/2529 J/PASJ/52/631 KOZASA T.: VII/60 KOZHURINA-PLATAIS V.: I/212 Kozhurina-Platais V.: I/277 I/288 J/A+A/461/509 J/AJ/113/1045 J/AJ/122/1486 J/AJ/124/601 Kozhurina-platais V.: J/AJ/126/2922 Kozlowski S.: J/MNRAS/370/435 Kozma C.: J/A+A/428/555 Kraakman H.: J/A+AS/101/87 Kraan-Korteweg R.: J/A+AS/141/123

Kraan-Korteweg R.C.: J/A+A/297/617 J/A+A/352/39 J/A+A/380/441 J/A+A/391/887 J/A+A/415/9

J/A+A/466/481 J/A+AS/127/463 J/AJ/124/1954 J/AJ/128/16 J/AJ/129/220 J/MNRAS/369/1131 KRAAN-KORTEWEG R.C.: VII/98A Kraemer K.E.: III/231 J/AJ/124/2990 J/AJ/126/1423 J/AJ/132/1445 J/ApJS/150/343 J/ApJS/151/299

V/114 Kraemer S.B.: J/AJ/132/321 Kraft R.P.: J/AJ/108/1722 J/AJ/112/545 J/AJ/113/279 J/ApJ/560/675 KRAICHEVA Z.: V/41 Krajci T.: J/other/IBVS/5592 Krajnovic D.: J/A+A/428/877 Kramer C.: J/A+A/426/119 J/A+AS/122/271 J/MNRAS/367/1609 Kramer M.: J/A+AS/147/195 J/MNRAS/328/17 J/MNRAS/335/275 J/MNRAS/342/1299 J/MNRAS/352/1439

J/MNRAS/353/1311 J/MNRAS/360/974 J/MNRAS/372/777 Kranich D.: J/A+A/390/39 J/A+A/406/L9 J/A+A/410/813 Krasberg M.: J/other/APh/26.282 Kraus A.: J/A+A/402/151 Kraus: J/A+A/364/391 Krause M.: J/A+A/472/785 Krause O.: J/A+A/422/39 J/A+A/453/923 J/A+A/466/1205 J/A+A/474/883 KRAUSHAAR W.L.: VII/43 Krauss M.I.: J/ApJ/589/509 Krautter J.: II/34A J/A+A/297/391 J/A+A/312/439 J/A+A/319/184 J/A+A/320/185 J/A+A/326/211

J/A+A/433/151 J/A+AS/123/329 J/A+AS/131/197 J/A+AS/132/341 J/ApJS/117/319 Krautter: J/A+AS/114/109 Kravtsov: J/A+A/469/529 Kravtsov V.: J/A+AS/109/479 J/A+AS/109/487 KRAVTSOV V.: J/A+AS/112/439 Kravtsov V.: J/A+AS/119/191 J/A+AS/125/1 J/AJ/112/2004 J/AJ/112/2020 J/AJ/114/189 J/AJ/116/2415 Kravtsov V.V.: J/PAZh/21/596 J/PAZh/22/269 J/PAZh/22/764 Kravtsova A.S.: J/AZh/78/963 Krawczynski H.: J/A+A/390/39 J/A+A/410/813 Kregenow J.M.: J/ApJ/560/675 KREINER J.M.: J/A+A/285/459 Kreiner J.M.: J/AcA/54/207 V/119 Krelowski J.: II/175 J/A+A/414/949 J/ApJ/634/451 J/MNRAS/358/563 KRENZIN R.: II/200 J/A+A/289/740 Kress K.: I/71 Kretchmer C.: J/AJ/127/180 J/ApJ/613/200 Kretke K.A.: J/ApJS/169/430 Kretschmar P.: J/A+A/395/129 Kreykenbohm I.: J/A+A/395/129 Kreysing H.C.: J/A+AS/114/465 Krichbaum: J/A+A/364/391 Krichbaum T.P.: J/A+A/402/151 J/A+A/476/L17 Kriek M.: J/AJ/134/1103 Kriessler J.R.: J/AJ/109/874 Krisciunas K.: J/A+A/336/920 J/AJ/130/2453 J/AJ/131/1648 J/ApJ/594/1 J/ApJ/602/571 Krishnamurthi A.: J/AJ/121/337 Kriss G.A.: J/A+A/369/57 J/AJ/104/495 J/ApJ/475/469 Krist J.: J/AJ/113/1001 J/ApJ/612/481 VII/149 Krist J.E.: J/AJ/128/1990 J/ApJ/612/202 J/ApJ/621/53 J/ApJS/150/1 Kristensen L.E.: J/A+A/469/459 Krivonos R.: J/A+A/475/775 Kriz S.: J/A+AS/125/75

Krockenberger M.: J/AcA/48/455 J/AJ/115/1016 J/AJ/115/1894 J/AJ/117/2810 J/AJ/118/2211

J/AJ/118/346 J/AJ/121/861 J/AJ/121/870 J/AJ/131/527 Krolik J.H.: J/A+A/369/57 J/AJ/125/1711 J/AJ/126/2125 J/AJ/128/1002 Kroll P.: J/A+A/310/384 J/A+A/454/295 Kron G.E.: II/8 J/AJ/110/1083 Kron R.: J/ApJS/159/41 Kron R.G.: J/A+AS/58/39 J/AJ/110/1083 J/ApJ/633/638 J/ApJS/109/45 J/ApJS/126/209 Kronberg P. P.: J/ApJS/124/285 KRONBERG P.P.: J/A+AS/105/385 Kronberg P.P.: J/A+AS/135/571 Kronberg P.P: J/ApJS/45/97 Kronberger M.: J/A+A/447/921 Krongold Y.: J/ApJ/639/37 Krpata J.: J/A+A/416/669 J/A+A/455/1037 Krticka J.: J/A+A/419/607 Kruegel E.: J/A+AS/118/47 Krughoff K.S.: J/AJ/124/1918 KRUIJSWIJK M.: II/200 Kruk J.W.: J/A+A/470/317 J/ApJS/146/1 Krumpe M.: J/A+A/466/41 J/A+A/474/473 Krusberg Z.A.C.: J/AJ/131/1565 Krywult J: J/A+AS/146/373

Krzeminski W.: I/244A J/A+A/424/1101 J/A+AS/112/407 J/A+AS/118/303 J/A+AS/122/471 J/A+AS/125/337 J/A+AS/125/343 J/A+AS/128/19 J/AJ/110/1141 J/AJ/110/2166 J/AJ/110/2206 J/AJ/112/2408 J/AJ/128/2239 J/MNRAS/365/110

Krzesinski J.: J/A+A/325/987 J/A+A/345/505 J/A+A/444/643 J/AJ/123/3409 J/AJ/126/1023

J/AJ/131/1674 J/AJ/132/676 J/ApJ/607/426 J/PASJ/55/771 Kubat J.: J/A+A/309/521 J/A+A/319/867 J/A+A/328/551 J/A+A/419/607

Kubiak M.: I/244A II/278 J/A+AS/112/407 J/A+AS/120/139 J/A+AS/122/471 J/A+AS/125/343 J/A+AS/128/19 J/AcA/48/147 J/AcA/48/175 J/AcA/48/563 J/AcA/49/223 J/AcA/49/319 J/AcA/49/437 J/AcA/49/521 J/AcA/50/1 J/AcA/51/175 J/AcA/51/221 J/AcA/51/259 J/AcA/51/317 J/AcA/52/129 J/AcA/52/143 J/AcA/52/217 J/AcA/53/1 J/AcA/53/93 J/AcA/54/347 J/AcA/56/253 J/AJ/110/1141 J/AJ/128/2815 J/ApJ/636/240 J/MNRAS/348/1439 J/MNRAS/356/331

Kuchar T.: J/AJ/126/1423 Kuchar T.A.: J/ApJ/488/224 Kuchinski L.E.: J/AJ/109/1131 J/AJ/109/1154 J/AJ/110/2844 J/ApJS/143/73 Kucinskas A.: J/A+A/442/281 J/ApJS/155/401 Kudritzki R.P.: J/A+A/306/9 J/A+A/321/531 Kudritzki R.-P.: J/A+A/437/467 Kudritzki R.P.: J/A+A/440/783 Kudritzki R.-P.: J/AJ/127/2723 J/AJ/128/1167 J/AJ/128/2815 J/ApJ/635/290 J/ApJS/161/154 Kudrya Y.N.: J/other/BSAO/50.5 VII/219 Kudrya Yu.N.: J/A+A/435/459 J/other/BSAO/57.5 J/PAZh/31/563 Kudryavtsev D.O.: J/other/Ap/46.234 Kudryavtsev S.M.: J/A+A/471/1069 Kudryavtseva N.A.: J/A+A/462/1081 J/PAZh/30/243 Kuehl D.: J/AN/323/484 Kuehn K.: J/other/APh/26.282 KUEHR H.: I/169 Kuehr H.: III/175 J/A+AS/105/211 J/A+AS/105/67 J/A+AS/115/1 J/A+AS/98/393 Kuehr M.: J/A+AS/115/11 Kuemmel M.: J/A+A/398/49 J/ApJS/115/185 Kuenzli M.: J/A+AS/127/277 Kuerster M.: IX/10A IX/11 J/A+A/311/211 KUESTNER F.: I/28 Kuhm B.: J/A+AS/109/537 Kuhn B.: J/A+A/319/67 J/A+AS/117/261 Kuhn O.P.: J/MNRAS/373/781 Kuhna K.M.: V/92A KUHR H.: VII/78 VIII/5

Kuijken K.: III/151 J/AJ/122/2205 J/AJ/125/1107 J/AJ/131/1891 J/ApJ/604/521 J/MNRAS/369/120

J/MNRAS/369/97 Kuimov K.: I/250 Kuiper L.: J/A+AS/143/145 Kuiper T.B.H.: J/ApJ/628/789 Kuipers J.A.: J/MNRAS/363/1019 Kuji S.: J/PASJ/52/631 V/69 Kukarkin B.V. et al: II/11 Kukarkin B.V. et al.: II/79 KUKARKIN B.V.: II/113 Kukarkin B.V.: II/12 II/13 II/140 II/69 II/70 Kukarkina N.P.: II/139B II/140 Kukula M. J.: J/ApJ/511/612 Kulagina N.B.: J/PAZh/24/93 Kulkarni S.R.: J/A+A/416/1037 J/AJ/121/2189 J/AJ/130/2866 J/ApJ/563/L49 J/ApJ/636/1020 Kulkarni V.: J/A+A/436/457 Kulkarni V.K.: J/A+A/468/1001 J/A+AS/82/41 Kulkarni V.P.: J/ApJ/571/665 Kulyk I.: J/A+A/383/724 KUMKOVA I.: I/169 Kumsiashvili M.I.: J/AZh/78/408 J/PAZh/25/764 Kun M.: J/A+A/272/235 Kun M: J/A+A/453/923 Kun M.: J/A+A/463/1017 J/ApJS/115/59 Kundu A.: J/AJ/131/866 J/ApJ/513/733 J/ApJ/586/814 Kuneva I.: J/A+AS/113/451 VII/190 VII/201 Kunkel M.: J/A+A/356/541 Kunkel W.E.: J/A+AS/122/463 J/AJ/119/2780 J/AJ/123/832 J/AJ/128/245 J/AJ/130/2677 J/ApJ/623/181 Kunszt P.Z.: J/AJ/122/503 KUNTH D.: J/A+A/285/1 Kunth D.: J/A+A/355/89 J/MNRAS/366/480 Kuntschner H.: J/A+A/434/53 J/A+A/454/423 J/A+A/478/83 J/MNRAS/323/839 J/MNRAS/326/473 Kuntz K.D.: J/ApJ/561/189 Kunze D.: J/MNRAS/351/1290 Kunzli M.: J/A+A/324/137 J/A+AS/122/51 Kunzst P.Z.: J/AJ/123/567 Kupka F.: J/A+A/328/349 J/A+A/405/1095 Kurabayashi T.: J/PASJ/56/295 Kuraszkiewicz J.K.: J/ApJS/143/257 J/ApJS/150/165 Kuratov K.S.: J/A+AS/147/5 Kurita M.: J/ApJ/596/1064 J/MNRAS/353/705 J/MNRAS/354/980 Kurk J.D.: J/A+A/431/793 J/A+A/461/823 Kurochkin N.E.: II/12 II/13 II/140 II/205 II/69 II/70 Kurochkin N.E. Medvedeva G.I.: II/139B KUROSU S.: VII/140 Kurpinska-Winiarska M.: J/A+A/367/865 Kurt V.G.: J/A+AS/114/341 Kurtanidze O.: J/A+A/376/775 Kurtanidze O.M.: J/A+A/402/151 J/other/Ap/31.714 Kurtev R.: J/A+A/400/917 J/A+A/407/919 Kurtz D .W.: J/A+A/374/1056 Kurtz D.W. et al.: II/146 Kurtz D.W.: J/A+A/455/303 J/ApJ/542/257 Kurtz M.: J/AJ/105/788

Kurtz M.J.: J/AJ/109/2368 J/AJ/116/1 J/AJ/122/2893 J/AJ/122/714 J/AJ/126/1362 J/AJ/126/2152

J/AJ/128/1078 J/AJ/130/1097 J/ApJ/518/69 J/ApJ/559/606 J/PASP/111/438 VII/200 Kurtz S.: J/ApJS/154/553 J/ApJS/155/149 J/ApJS/91/659 KURUCZ R.L.: J/A+A/281/817 Kurucz R.L.: J/A+A/454/333 J/A+AS/122/51 KURUCZ R.L.: J/ApJ/416/787 Kurucz R.L.: VI/10 KURUCZ R.L.: VI/33 VI/39 VI/77 Kurucz R.L.: VI/77 KURUCZ R.L.: VI/78 Kus A.: J/PASJ/52/997 Kus A.J.: J/A+A/474/1093 J/A+AS/143/269 Kus: J/A+A/413/233 Kusakin A.V.: J/AZh/78/408 J/PAZh/28/764 Kuschnig R.: J/A+A/373/625 Kuss C.: II/202 J/A+A/304/415 KUSTNER F.: I/84 Kutkin A.M.: J/PAZh/31/332 Kutner M.L.: J/A+A/276/25 Kuulkers E.: J/A+A/242/401 J/A+A/467/585 Kuwashima F.: J/ApJ/382/636 Kuzio de Naray R.: J/ApJ/577/31 Kuzmin A.: I/208 I/250 Kuz’min A.D.: J/AZh/79/501 Kuzmin A.V.: III/182 Kuznetsov A.V.: IX/36 Kuznetsov V.I.: III/230 Kuznetsova Eh.F.: V/124 Kuznietsova J.: J/A+A/430/255 Kwiatkowski T.: J/A+A/465/331 Kwok P.W.: J/ApJS/94/551 Kwok S.: J/A+A/368/1006 J/AJ/127/3121 J/ApJS/112/557 J/ApJS/152/201 Kwon S.-G.: J/A+A/404/621 J/A+A/405/231

L.: J/A+A/413/233 J/A+A/421/937 J/A+A/445/939 J/MNRAS/296/839 La Barbera F.: J/A+A/389/787 J/A+A/424/79 J/A+A/425/797 J/MNRAS/358/1116 La Franca F.: J/A+A/321/123 J/A+A/370/900 J/A+A/409/79 J/A+A/457/501 J/A+A/466/31 LA FRANCA F.: J/A+AS/112/347 La Franca F.: J/A+AS/140/351 J/AJ/127/3075 J/ApJ/570/100 J/MNRAS/327/771 J/MNRAS/342/575 La Palombara N.: J/A+A/458/245 Laakso T.: J/A+A/350/513 Laan K.: J/A+A/352/228 Labat J.: J/A+AS/102/607 Labbe I.: J/AJ/122/2205 J/AJ/125/1107 J/AJ/131/1891 J/ApJ/604/521 Labhardt L.: J/ApJ/522/802 J/ApJ/551/973 Labiano A.: J/A+A/463/97 LABONTE A.E.: J/AJ/102/395 Lacasse M.G.: J/ApJ/620/961 Lacerda P.: J/A+A/407/1139 Lacey C.: J/MNRAS/372/1621 Lacey J.D.: J/AJ/74/366 Lachaume R.: J/A+A/348/897 Lachieze-Rey M.: J/A+A/303/137 LACHIEZE-REY M.: J/A+AS/109/447 Laclare F.: J/A+AS/135/227 Lacombe F.: J/A+A/363/62 Lacroute P.: I/2 LACROUTE P.: I/21A I/21A I/22A I/22A Lacy C.H.: J/ApJ/640/1018 Lacy C.H.S.: J/AJ/114/326 J/AJ/119/1389 J/AJ/128/1324 J/AJ/128/1340 J/AJ/128/1840 J/AJ/130/2838 Lacy M.: J/A+AS/108/79 J/AJ/131/250 J/ApJ/607/60 J/ApJS/135/227 J/MNRAS/276/614 J/MNRAS/307/293

J/MNRAS/344/283 Lada C.: J/A+A/462/L17 J/ApJ/638/897 Lada C.J.: J/AJ/111/1964 J/AJ/113/1788 J/AJ/128/1254 J/AJ/130/1145 J/AJ/131/1574 J/AJ/131/527 LADA C.J.: J/ApJ/397/520 Lada C.J.: J/ApJ/573/366 J/ApJ/575/354 J/ApJ/593/1093 J/ApJS/160/319 Lada E.A.: J/AJ/111/1964 J/AJ/113/1788 J/AJ/128/1254 J/ApJ/573/366 J/ApJ/593/1093 J/ApJS/112/109 Ladd E.F.: J/A+A/440/151 J/A+A/468/1009 J/MNRAS/367/1609 Ladjal D.: VI/125 Lagache G.: J/A+A/372/173 J/A+A/372/364 Lagadec E.: J/MNRAS/376/313 Lagage P.-O.: J/A+A/473/369 Lagerkvist C.-I: J/A+AS/113/115 Lagerkvist C.-I.: J/A+AS/115/295 VII/108C Lagerros J.: J/A+AS/113/115 Lagerros J.S.V.: J/A+A/338/340 Laget M.: J/A+A/250/370 J/A+A/303/661 J/A+A/324/915 Lago M.T.V.T.: J/A+A/466/137 J/A+A/476/199 Lagos P.: J/A+A/476/89 LAGRANGE-HENRI A.M.: J/A+A/299/557 Lagrange-Henri A.-M.: J/A+AS/136/429 Lahav O.: J/MNRAS/274/1107 J/MNRAS/277/125 J/MNRAS/329/227 J/MNRAS/329/87 J/MNRAS/333/100

J/MNRAS/350/1485 J/MNRAS/352/939 J/MNRAS/375/68 VII/149 Lahteenmaki A.: J/A+A/469/451 J/A+AS/116/157 J/A+AS/132/305 J/A+AS/143/357 J/AJ/133/1947 Lahtenmaki A.: J/A+A/427/769 Lahuis F.: J/A+A/390/1033 LAHULLA F.: J/A+AS/114/489 J/A+AS/94/519 Lahulla J.: J/A+AS/123/219 Lahulla J.F.: J/A+A/416/811 J/A+AS/113/325 J/A+AS/126/73 J/AJ/127/1181 Lai D.K.: J/AJ/128/2402 Lai O.: J/A+A/412/447 Laidler V.G.: J/ApJ/613/200 Laine S.: J/A+AS/116/157 J/AJ/125/478 J/AJ/131/250 J/MNRAS/317/234 Lainela M.: J/A+A/427/769 J/A+AS/116/157 J/A+AS/132/305 J/A+AS/143/357 Lainey V.: J/A+A/427/371 Laing J.D.: J/other/SAAOC/14.11 Laing R.A.: J/A+A/472/805 J/MNRAS/204/151 Laird E.S.: J/ApJ/642/126 J/ApJS/156/35 J/MNRAS/350/785 J/MNRAS/356/568 J/MNRAS/371/221 LAIRD J.B.: J/AJ/107/2240 Laird J.B.: J/AJ/124/1144 J/AJ/131/2274 Lakatos B.: J/A+A/430/1049 J/A+A/445/617 LAKAYE J.M.: VI/8 Lal D.V.: J/A+A/425/99 Lallement R.: J/A+A/346/785 J/A+A/411/447 Lamareille F.: J/A+A/448/893 J/A+A/457/79 J/MNRAS/362/1143 Lamb D.: J/AJ/130/968 J/other/GCN/1070.1 Lamb D.Q.: J/AJ/121/2308 J/AJ/121/2331 J/AJ/122/503 J/AJ/123/567 J/AJ/124/1651 J/AJ/125/1817

J/AJ/126/1023 J/AJ/126/2579 J/ApJ/588/824 J/ApJ/607/426 J/ApJS/145/245 Lambas D.G.: J/AJ/110/1032 J/AJ/129/1237 J/AJ/130/2012

Lambert D.L.: J/A+A/275/101 J/A+A/327/587 J/A+A/336/263 J/A+A/352/555 J/A+A/353/287 J/A+A/420/183 J/A+AS/102/603 J/AJ/114/341 J/ApJ/479/427 J/ApJ/527/879 J/ApJ/638/1018 J/MNRAS/340/304 J/MNRAS/349/757 J/MNRAS/367/1329 J/PASP/115/22

Lambert: J/A+A/465/271 Lambert S.: J/A+A/400/1145 Lamberts A.: VI/30A Lamer G.: J/A+A/311/384 J/A+A/428/383 J/A+A/466/41 J/A+A/469/27 J/A+A/470/821 J/A+A/476/1191

Lamers H.J.G.L.M.: J/A+A/306/924 J/A+A/307/869 J/A+A/310/893 J/A+A/397/473 J/A+A/420/571

J/A+A/428/149 J/A+AS/106/141 J/A+AS/98/505 J/ApJS/95/163 Laming J.M.: J/A+AS/126/281 Lamm M.H.: J/A+A/417/557 J/A+A/430/1005 LAMONTAGNE R.: J/AJ/105/35 Lamontagne R.: J/AJ/109/817 J/AJ/112/2408 Lampeitl H.: J/A+A/390/39 J/A+A/406/L9 J/A+A/410/813 Lampens P.: II/242 J/A+A/367/865 J/A+A/374/132 J/A+A/418/249 J/A+A/422/1023 J/A+A/437/705

J/A+A/438/201 J/A+A/464/641 J/A+AS/112/291 LAMPENS P.: J/A+AS/113/483 Lampens P.: J/MNRAS/259/563 J/MNRAS/356/545 Lampton M.: II/203A II/203A IX/35 IX/35 J/ApJS/111/143 J/ApJS/93/569 J/ApJS/96/461 J/MNRAS/313/641

J/MNRAS/324/580 J/MNRAS/332/441 Lamy H.: J/A+A/340/371 J/A+A/427/107 J/A+A/433/757 J/A+A/441/915 J/A+AS/142/451 Lamzin S.A.: J/A+AS/114/341 Lanchares V.: J/A+A/416/811 Lancon A.: III/196 III/232 J/A+A/393/167 J/A+A/402/425 J/A+A/405/531 J/A+AS/146/217 VI/125 Landaberry S.J.C.: J/A+A/333/658 LANDAU R.: J/AJ/102/395 Landecker T.L.: J/A+A/241/551 J/A+A/448/411 LANDECKER T.L.: J/A+AS/105/385 Landi E.: J/A+A/370/1071 J/A+A/444/305 J/A+AS/125/149 J/ApJ/543/1016 Landi O.: J/A+A/434/365 LANDOLT A.U.: II/118 Landolt A.U.: II/183A LANDOLT A.U.: VI/19 VI/19 Landsman H.: J/other/APh/26.282 Landsman W.B.: J/AJ/108/1350 J/AJ/122/2569 J/AJ/131/866 LANDSMAN W.B.: J/ApJ/408/305 Landsman W.B.: J/ApJ/562/368 J/ApJS/104/287

Landstreet J.D.: J/A+A/276/142 J/A+A/361/991 J/A+A/374/265 J/A+A/458/569 J/A+A/475/1053

J/A+AS/123/353 J/ApJS/82/117 J/MNRAS/376/361 Landt H.: J/A+A/470/787 J/MNRAS/323/757 J/MNRAS/351/83 Landy S.D.: J/ApJS/130/237 VII/203 Lane A.P.: J/ApJ/517/282 Lane B.F.: J/AJ/130/2866 J/ApJ/636/1020 Lane W.M.: VIII/79 Laney C.D.: J/AJ/132/111 J/MNRAS/224/61 Lang C.: J/MNRAS/371/38 Lang J.: J/A+A/432/1063 Lang R.: J/other/APh/26.282 Lang R.H.: J/MNRAS/342/738 J/MNRAS/346/787 J/MNRAS/355/1303 Lang W.J.: J/A+A/357/1001 Langan S.: J/AJ/128/1748 J/AJ/129/2007 J/AJ/130/832 Lange A.: J/ApJ/638/293 Langer G.E.: J/AJ/112/545 J/AJ/113/279 J/ApJ/518/262 Langer N.: J/A+A/305/960 J/A+A/437/467 Langston G.: J/AJ/119/2801 J/ApJS/75/801 J/PASJ/52/997 VIII/52B Langston G.I.: VIII/52B Lanning H.H.: II/212A II/231 II/257 J/PASP/118/1639 J/PASP/96/707 Lanoix P.: J/A+AS/135/133

Lanteri L.: J/A+A/395/11 J/A+A/402/151 J/A+AS/121/119 J/A+AS/127/445 J/A+AS/129/577

J/A+AS/130/109 J/A+AS/132/361 J/A+AS/144/481 LANZ T.: II/116 Lanz T.: J/A+A/323/881 J/A+A/348/897 J/A+A/434/1125 J/A+A/436/687 J/A+A/466/277 J/A+AS/123/353

J/ApJ/562/368 J/other/Ap+SS/163.275 Lanza A.F.: J/A+A/429/309 Lanzafame A.C.: J/A+A/432/1063 Lanzetta K.M.: J/AJ/123/2206 J/ApJ/498/77 J/ApJ/513/34 J/ApJ/538/493 J/ApJ/559/654 J/ApJS/135/41 Laor A.: J/ApJ/632/799 J/ApJS/105/369 Lapaev K.A.: J/AZh/77/21

Lapasset E.: J/A+A/314/448 J/A+A/361/945 J/A+A/399/543 J/A+A/463/789 J/A+AS/109/375

J/A+AS/128/131 J/PASP/108/332 J/PASP/108/338 Lapasset: J/A+A/409/541 LAPIERRE G.: J/A+A/287/591 LAPINOV A.: J/A+A/288/601 Lapinov A.V.: J/PAZh/33/143 Lara L.: J/A+A/370/409 Large M.I.: VIII/16 VIII/2

Lari C.: J/A+A/427/23 J/A+A/427/773 J/AJ/127/3075 J/MNRAS/325/1173 J/MNRAS/351/1290

J/MNRAS/358/333 J/MNRAS/358/397 J/MNRAS/371/1891 Larianov M.G.: VIII/49 VIII/49 Larink J.: I/16 Larionov M.: J/A+AS/114/341 Larionov M.G.: J/AZh/84/387 Larionov V.: J/A+A/462/1081 Larionov V.M.: J/A+A/385/55 J/PAZh/30/243 Larionova E.G.: J/A+A/385/55 J/PAZh/30/243 Larios G.R.: J/MNRAS/357/732 Larkin J.E.: J/ApJ/581/258 Larosa T.N.: J/AJ/128/1646 Larrarte J.J.: VIII/75 Larsen J.A.: II/254 Larsen S.: J/AJ/127/2114

Larsen S.S.: J/A+A/364/455 J/A+A/402/509 J/A+A/416/537 J/A+A/443/413 J/A+A/459/391

J/A+A/464/495 J/A+AS/122/559 J/A+AS/139/393 J/MNRAS/327/1116 Larson A.M.: III/185 Larson K.A.: J/ApJ/623/897 Larson S.B.: J/ApJ/491/93 Larsson A.: J/other/IBVS/5633 Larsson S.: J/A+A/452/975 Lasenby A.: J/MNRAS/341/1066 LASKAR J.: VI/63 VI/66 LASKARIDES P.G.: VI/55 Lasker B.: J/A+AS/137/305 Lasker B.M.: I/220 I/254 I/255 II/143A J/A+AS/130/359 Lasserre T.: J/A+A/377/945 J/A+A/389/149 LASTENNET E.: J/A+AS/108/611 Latham D.: J/A+A/473/829

Latham D.W.: J/A+A/339/423 J/A+A/375/L27 J/A+A/399/105 J/A+A/415/531 J/A+AS/126/21

J/AJ/105/1637 LATHAM D.W.: J/AJ/107/2240 Latham D.W.: J/AJ/108/1987 J/AJ/116/1 J/AJ/119/1389 J/AJ/124/1144 J/AJ/129/2294 J/AJ/131/2274

J/ApJ/466/415 J/PASP/106/281 J/PASP/118/1656 Latini G.: J/A+A/327/61 J/A+AS/121/119 Latorre A.: J/A+AS/146/103 Lattanzi M.G.: I/176 I/202 J/A+A/341/L79 J/A+A/448/579 J/A+AS/121/119 J/A+AS/130/359 Latter W.: J/ApJS/154/155 Latter W.B.: J/AJ/131/250 J/ApJS/124/195 Lauberts A.: VII/115 VII/34C Laubscher B.E.: J/AJ/127/1883 Lauer T.R.: J/AJ/125/478 J/AJ/129/2138 J/ApJ/425/418 Lauer T.R: J/ApJ/440/28 Lauer T.R.: J/ApJ/517/650 J/ApJ/579/93 Laugalys V.: J/BaltA/11/219 J/BaltA/11/231 J/BaltA/12/323 J/BaltA/13/1 J/BaltA/15/483Short Lauger S.: J/ApJ/619/L95 Laughlin G.: J/ApJ/634/625 J/ApJ/637/1102 J/ApJS/169/430 Launay F.: J/A+AS/128/361 Launhardt R.: J/AJ/129/1049 Laureijs R.J.: J/A+A/365/545 Laurent B.: J/A+A/303/137 LAURENT B.: J/A+AS/109/447 Laurent G.: J/ApJ/638/293 Laurent-Muehleisen S.A.: J/A+A/323/739 J/A+A/356/445 J/A+AS/122/235 J/AJ/131/1948 J/ApJ/525/127

J/ApJS/126/133 J/ApJS/135/227 VIII/71 Laurikainen E.: J/A+AS/141/103 J/MNRAS/355/1251 Lauroesch J. T.: J/ApJS/122/355 Lautsen S.: I/99 Law D.R.: J/AJ/128/245 Lawler J.E.: J/ApJ/617/1091 J/ApJ/645/613 J/ApJS/162/227 Lawler S.M.: J/ApJ/636/1098 Lawrence A.: J/AJ/129/2074 J/MNRAS/253/485 J/MNRAS/274/1194 J/MNRAS/279/459 J/MNRAS/302/222

J/MNRAS/305/297 J/MNRAS/308/897 J/MNRAS/343/293 J/MNRAS/344/887 J/MNRAS/351/1290 Lawrence C.R.: J/ApJS/51/67 VIII/52B Laws C.: J/MNRAS/378/1141 Lawson W.A.: J/ApJ/614/267 Laychak M.B.: J/ApJ/577/31 J/ApJ/614/167 Layden A.: J/AJ/113/264 LAYDEN A.C.: III/176 Layden A.C.: J/AJ/108/1016 J/AJ/109/1086 J/AJ/110/2288 J/AJ/112/2408 J/AJ/115/193 J/AJ/117/1313

J/AJ/117/2329 J/AJ/119/1760 J/ApJ/632/266 J/PASP/109/524 J/PASP/109/977 Lazaro C.: J/A+A/375/527 Lazauskaite R.: J/BaltA/5/1 Lazendic J.S.: J/A+AS/127/119 Lazio T.J.W.: J/AJ/128/1646 J/AJ/132/242 J/ApJ/612/511 J/ApJS/118/201 J/ApJS/136/265 J/ApJS/150/417

J/ApJS/155/89 J/ApJS/159/242 J/ApJS/165/439 VIII/79 Lazorenko P.: J/A+A/438/377 Lazrek M.: J/A+A/390/717 LAZZARIN M.: J/A+AS/112/347 Lazzaro D.: J/A+AS/134/463 Lazzati D.: IX/34 Le Bail K.: J/A+A/375/661 LE BERTRE T.: II/177

Le Bertre T.: J/A+A/227/82 J/A+A/314/896 J/A+A/324/1059 J/A+A/334/609 J/A+A/361/641

J/A+A/363/62 J/A+A/376/997 J/A+A/387/1 J/A+A/403/943 J/A+A/422/129 J/A+A/475/559 J/A+AS/97/729 LE BERTRE T.: J/A+AS/99/31 Le Bertre T.: J/PASP/117/199 Le Borgne D.: III/251 Le Borgne J.-F.: III/232 J/A+A/448/893 Le Borgne J.F.: J/A+AS/95/87 Le Borgne J.-F.: J/MNRAS/330/1 Le Borgne J.F.: J/other/IBVS/5767 Le Brun V.: J/A+A/403/857 J/A+A/428/1043 J/A+A/457/79 Le Campion J.-F.: I/272 Le Campion J.F.: I/303 J/A+A/383/1054 Le Campion J.-F.: J/A+A/409/361 J/A+AS/119/307 Le Campion J.F.: J/A+AS/125/399 J/A+AS/136/531 LE COARER E.: J/A+A/280/365 Le Contel D.: J/A+A/304/406 J/A+A/448/697 Le Contel J.M.: J/A+A/304/406 J/A+A/312/879 Le Contel J.-M.: J/A+A/426/247 J/A+A/448/697 Le Dourneuf M.: J/A+A/394/753 J/A+AS/119/99 J/A+AS/120/361 Le Du Y.: J/A+A/421/509 Le Fevre J.-P.: J/MNRAS/382/279 Le Fevre O. et al.: III/250 Le Fevre O.: J/A+A/310/31 J/A+A/310/8 J/A+A/403/857 J/A+A/428/1043 J/A+A/439/413 J/A+A/457/79

J/A+A/463/519 J/A+A/467/73 J/A+A/468/937 J/A+A/474/473 J/A+AS/111/265 J/A+AS/122/409 J/AJ/128/1974 J/ApJ/464/79 J/ApJ/481/49 J/ApJ/499/112 J/ApJ/517/148 J/MNRAS/382/279 VII/225A

Le Floch B.: J/A+AS/102/611 Le Floc’h E.: J/AJ/132/231 J/ApJ/634/L1 J/ApJ/640/167 J/ApJS/154/155 J/MNRAS/380/986 Le Floch J.C.: J/A+AS/125/399 Le Guillou L.: J/A+A/389/149 J/A+A/454/185 J/A+A/463/243 Le Guyader C.: J/A+AS/136/445 Le Mignant D.: J/AJ/127/3121 Le Poole R.S.: J/A+A/360/472 Le Roy R.J.: J/A+AS/114/175 Le Sidaner P.: J/A+A/314/896 Le Squeren A.M.: J/A+A/277/453 LE SQUEREN A.M.: J/A+A/297/494 Le Squeren A.M.: J/A+AS/111/237 Le Teuff Y.H.: J/A+A/357/572 Leahy D.A.: J/A+AS/115/345 Leahy J.P.: J/MNRAS/341/1066 Leao I.: J/A+A/466/L1 Learner R.C.M.: J/ApJS/94/221 Lebedev V.S.: VII/77 VII/82A LEBEDEV V.S.: VII/83 Lebre A.: J/A+A/423/353 J/A+A/475/1053 Lebreton Y.: J/A+A/331/81 Lebrun F.: J/ApJ/607/L33 J/ApJ/636/765 J/ApJS/170/175 Lebrun J.C.: J/A+A/478/197 Lebzelter T.: J/A+A/375/527 J/A+A/397/997 Lecacheux J.: J/A+AS/125/399 Lecampion J.F.: J/A+A/438/769 LECAVELIER DES ETANGS A.: J/A+A/299/557 Lecavelier des Etangs A.: J/A+AS/145/275 Leccia S.: J/A+A/459/557 Leckrone D.S.: J/AJ/116/941 Lecureur A.: J/A+A/465/799 Lederle C.: J/A+A/413/1037 Lederle T.: I/149A I/175 Ledlow M.: J/ApJ/611/821 Ledlow M.J.: J/AJ/109/14 J/AJ/109/853 J/AJ/110/1959 J/AJ/118/2014 J/AJ/120/2269 J/AJ/123/3018

J/AJ/126/119 J/AJ/126/2740 J/AJ/129/26 J/AJ/129/31 J/AJ/130/47 J/ApJ/581/844 J/ApJ/613/841 J/ApJS/146/267

Ledoux C.: J/A+A/370/426 J/A+A/436/457 J/A+A/473/791 J/A+AS/126/311 J/MNRAS/367/1478 Ledovskaya I.V.: J/other/KFNT/15.483 Lee B.: J/AJ/119/1901 Lee B.C.: J/AJ/126/2579 J/AJ/127/1811 J/AJ/127/1883 J/AJ/127/2436 J/AJ/129/2062 Lee B.-C.: J/PASP/118/636 Lee B.-W.: J/AJ/123/290 Lee C.-U.: J/A+A/404/621 J/A+A/450/39 Lee C.W.: J/ApJ/526/788 J/ApJS/123/233 Lee C.-W.: J/ApJS/165/360 Lee D.-J.: J/A+A/404/621 Lee H.: J/ApJS/153/429 Lee H.-C.: J/ApJS/160/176 Lee H.-H.: J/A+AS/119/111 Lee H.-W.: J/AJ/128/846 J/MNRAS/310/982 Lee J.C.: J/ApJS/165/307 Lee J.H.: J/AJ/126/2840 Lee J.-K.: J/A+A/410/597 J/A+A/437/467 Lee J.K: J/A+A/457/265 Lee J.T.: I/238A Lee J.W.: J/A+A/405/231 Lee J.-W.: J/AJ/117/2868 J/AJ/129/656 Lee J.W.: J/MNRAS/379/1665 Lee K.: J/ApJ/613/200 Lee K.H.: J/MNRAS/375/115 LEE K.M.: J/AJ/107/679 Lee K.M.: J/AJ/128/1748 J/AJ/129/2007 J/AJ/130/832 Lee M.G: J/A+A/454/717 LEE M.G.: J/AJ/106/493

Lee M.G.: J/AJ/106/66 J/AJ/106/964 J/AJ/112/1438 J/AJ/119/777 J/AJ/126/2840 J/ApJ/324/172

J/ApJ/559/812 J/MNRAS/375/115 Lee O.J.: II/68 II/68 II/68 Lee S.-G.: J/AJ/123/290 J/AJ/126/2840 Lee S.-W.: J/AJ/114/2644 J/AJ/115/734 J/MNRAS/310/982 Lee Y.P.: J/ApJ/569/689

Lee Y.-W.: J/AJ/116/1775 J/AJ/119/1824 J/AJ/122/3219 J/AJ/126/2840 J/AJ/127/958 J/AJ/130/825

J/ApJ/619/L95 J/ApJS/164/38 Leech K.: J/A+A/357/839 J/A+A/365/545 J/A+A/407/791 J/MNRAS/253/485 Leedjarv L.: J/A+A/415/273 Lefever K.: J/A+A/455/259 J/A+A/463/243 J/A+A/465/965 Lefloch B.: J/ApJ/607/904 J/ApJ/643/965 Leger R.F.: J/AJ/121/2308 J/AJ/121/2331 J/AJ/123/567 Leggett S.K.: J/AJ/127/3553 J/AJ/131/2722 J/ApJ/637/1067 J/ApJS/133/413 J/MNRAS/373/781 VI/49 LEGGETT S.K.: VII/169 Legrand E.: V/53A Legrand F.: J/A+A/376/775 Lehar J.: J/AJ/130/1977 J/ApJ/510/64 VIII/52B VIII/52B Lehmann H.: J/A+A/367/236 J/A+A/392/851 J/A+A/408/611 J/A+A/413/293 J/A+A/457/1033

Lehmann I.: J/A+A/320/776 J/A+A/371/833 J/A+A/393/425 J/A+A/398/901 J/A+A/428/383

J/A+A/444/79 J/ApJS/155/271 J/PASJ/53/445 Lehmer B.D.: J/ApJS/161/21 Lehner M.: J/ApJ/542/257 J/ApJ/598/597 Lehner M.J.: J/AJ/114/326 J/AJ/127/334 J/ApJ/552/289 J/ApJ/562/337 Lehnert M.: J/A+A/465/1099 Lehnert M.D.: J/A+A/406/593 J/A+A/443/41 J/ApJ/606/664 Lehtinen K.: J/A+A/405/999 J/A+A/421/1087 J/A+A/463/1029 J/A+AS/112/99 J/A+AS/132/211 Lei C.M.: J/other/PPMtO/14.185 Leibowitz E.: J/A+A/459/557 LEIBUNDGUT B.: II/189 Leibundgut B.: J/A+A/378/556 J/A+A/428/555 J/A+A/429/559 LEIBUNDGUT B.: J/AJ/107/1022 Leibundgut B.: J/AJ/130/2453 J/AJ/131/1648 J/ApJ/594/1 J/ApJ/602/571 J/ApJ/607/665 J/ApJ/613/200 Leich H.: J/other/APh/26.282 Leighly K.M.: J/AJ/127/156 J/ApJS/115/185 Leighton R.B.: II/2B Leinert C.: J/A+A/314/846 J/A+A/325/159 J/A+A/356/541 J/A+A/369/249 J/A+A/376/982 J/A+A/405/999

J/A+AS/112/99 J/A+AS/127/1 LEIR A. YOUNGER F.: III/29 Leir A.A.: VII/87A Leisawitz D.T.: VII/106 Leister N.V.: J/A+AS/137/269 J/MNRAS/371/252 Leistra A.: J/AJ/130/1719 Leisy P.: J/A+A/456/451 J/A+AS/116/95 J/A+AS/121/407 Leitherer C.: J/AJ/117/1708 J/ApJ/619/755 J/ApJ/633/871 J/ApJS/144/21 J/ApJS/99/173 J/MNRAS/357/945 Lejeune T.: J/A+AS/125/229 J/A+AS/130/65 VI/102 Lelievre M.: J/AJ/120/1306 Lemaire J.L.: J/A+A/453/517 J/A+A/466/949 J/A+A/469/459 Lemaire P.: J/A+A/375/591 J/A+AS/126/281 Lemke D.: J/A+A/359/865 J/A+A/372/364 J/A+A/401/873 J/A+A/405/999 J/A+A/422/39 J/A+A/466/1205

J/A+A/474/883 J/MNRAS/302/222 J/MNRAS/305/297 J/MNRAS/351/1290 Lemke M.: J/A+A/400/939 J/A+AS/122/285 Lemke R.: J/A+A/376/1123 J/A+AS/118/47 Lemon D.J.: VII/240 Lemonon L.: J/A+A/363/440 J/A+A/392/795 Lenhardt H.: I/264 I/286

Lennon D.J.: J/A+A/321/531 J/A+A/334/987 J/A+A/367/86 J/A+A/434/1125 J/A+A/436/687 J/A+A/437/467 J/A+A/456/623 J/A+A/457/265 J/A+A/466/277 J/A+A/471/625 J/A+AS/103/273

Lenorzer A.: J/A+A/384/473 J/A+A/390/1033 Leon S.: J/A+A/411/391 J/A+A/436/443 J/A+A/449/937 J/A+A/462/507 J/A+A/464/L21 J/A+A/470/505

J/A+A/472/121 J/A+AS/122/271 Leonard D. C.: J/ApJ/510/659 Leonard D.C.: J/ApJS/115/185 Leone F.: J/A+A/464/939 LEONE F.: J/A+AS/108/595 Lepage K.: J/MNRAS/372/1621 Lepardo A.: J/A+A/323/113 J/A+AS/116/1 Lepeintre F.: J/A+A/303/137 Lepine J.R.D.: B/ocl J/A+A/227/82 LEPINE J.R.D.: J/A+A/299/453

Lepine J.R.D.: J/A+A/361/1143 J/A+A/376/441 J/A+A/381/32 J/A+A/388/168 J/A+A/392/491

J/A+A/404/913 J/A+A/413/159 LEPINE J.R.D.: J/A+AS/99/31

Lepine S.: I/298 J/AJ/124/1190 J/AJ/125/1598 J/AJ/130/1247 J/AJ/130/1680 J/PASP/118/1238

J/PASP/118/1639 Lequeux J.: J/A+A/276/25 J/A+A/327/522 J/A+A/406/867 J/A+A/411/381 J/A+AS/102/251 J/AJ/77/405

VII/15 Lercher G.: J/A+AS/117/369 Leroy J.L.: J/A+A/346/955 J/A+AS/101/551 J/A+AS/114/79 Lesh J.R.: II/96 III/23B LESH J.R.: V/10 V/9 Lespine Y.: J/A+A/317/416 Lesquoy E.: II/279 J/A+A/303/137 J/A+A/377/945 J/A+A/389/149 J/A+A/454/185 LESQUOY E.: J/A+AS/109/447 Lesser M.P.: J/AJ/128/1990 J/ApJ/612/202 J/ApJ/621/53 J/ApJS/150/1 Lestrade J.P.: IX/20A Letarte B.: J/A+A/410/795 Letarte B: J/A+A/424/125 Letarte B.: J/A+A/459/423 J/AJ/123/832 J/AJ/125/1298 J/AJ/125/3037 Letawe G.: J/A+A/391/481 J/A+A/478/321 Letawe Y.: J/A+A/478/321 Leto G.: J/A+A/327/1114 LEUENHAGEN U.: J/A+AS/103/445 Leurini S.: J/A+A/466/1065 J/A+A/475/925 Leuthold M.: J/other/APh/26.282 Levasseur-Regourd A.-C.: J/A+AS/127/1 Levato H.: B/sb9 III/204 III/216 III/249 III/58 J/A+A/408/305 J/A+AS/118/231 J/ApJ/573/359 V/122 Levay Z.: J/AJ/112/1335 J/AJ/120/2747 J/ApJS/150/1 Levenhagen R.S.: J/MNRAS/371/252 Levenson N.A.: J/ApJ/581/258 Levin Y.: J/ApJ/643/1011 Levine A.: J/other/GCN/1070.1 Levine S.E.: I/284 I/297 J/AJ/127/3034 J/AJ/131/582 Levison H.F.: J/AJ/116/1998 Lewin W.: J/ApJS/156/217 Lewin W.H.G.: II/208 J/A+A/278/36 J/A+A/373/63 J/A+AS/114/215 J/A+AS/126/401 J/ApJS/134/385 Lewis A.D.: J/ApJ/566/744 Lewis B.M.: J/A+AS/111/237 J/AJ/127/501 J/ApJS/109/489 J/ApJS/89/189 J/ApJS/93/549 Lewis C.: J/AJ/130/794 J/PASJ/52/997 Lewis D.W.: VII/171 Lewis G.F.: J/MNRAS/339/701 J/MNRAS/380/281 Lewis I.: J/MNRAS/329/227 J/MNRAS/329/87 J/MNRAS/333/100 J/MNRAS/350/1485 J/MNRAS/352/939

J/MNRAS/362/1143 Lewis I.J.: VII/223 VII/241 Lewis J.: II/203A II/203A IX/35 IX/35 J/ApJS/93/569 Lewis J.W.: J/AJ/107/751 Lewis M.S.: J/AJ/131/2538 Lewitus E.: J/AJ/132/891 Leyder J.C.: J/A+A/470/1009 Li D.: J/A+AS/118/239 J/A+AS/132/233 Li F.L.: J/A+A/357/225 Li J.: J/A+AS/120/201 J/AJ/127/3587 J/ApJS/160/319 Li J.Z.: J/A+AS/132/173 LI K.H.: J/A+AS/106/361 Li K.H.: J/A+AS/132/83 J/A+AS/136/455 J/ApJ/522/846 Li L.: J/MNRAS/370/1181 Li L.-X.: J/AJ/126/1720 Li Q.-K.: J/AJ/123/578

Li W.: J/A+A/428/555 J/AJ/130/2453 J/AJ/131/1648 J/AJ/131/2245 J/ApJ/594/1 J/ApJ/602/571

J/ApJ/607/665 J/ApJ/627/579 J/ApJ/636/1002 J/PASP/114/820 J/PASP/115/1280 J/PASP/117/773 Li X.: J/A+AS/118/239 Li X.C.: J/other/AcApS/16.103 Li Y.: J/A+AS/118/239 J/AJ/112/628 J/AJ/117/2757 Li Z.: J/A+A/471/795 J/ApJ/645/488 J/MNRAS/371/147 J/MNRAS/376/960 Li Z.P.: J/A+A/468/637 Li Z.-P.: J/A+AS/115/469 Li Z.P.: J/A+AS/136/515 Li Z.W.: J/ApJS/126/461 Li Z.-Y.: J/A+AS/131/115 Liang E.-W.: J/PASJ/55/947 Liang H.: J/A+A/363/440 J/A+A/456/791 Liang S.-G.: J/PASJ/52/997 Liang Y.C.: J/A+A/397/257 J/A+A/423/867 J/A+A/430/115 J/A+A/462/535 J/A+A/465/1099 Liao T.X.: J/ApJ/633/174 Libbrecht C.: J/MNRAS/382/279 Libero V.: J/A+A/446/949 Libich J.: J/A+A/446/583 J/A+A/455/1037 Licandro J.: J/A+AS/146/73 Lichau H.: J/A+A/398/L11 Lichti G.G.: J/A+A/438/1175 J/A+AS/143/145 LICK OBS.: VI/31 Lidman C.: J/A+A/432/381 J/A+A/437/789 J/MNRAS/306/708

Liebert J.: II/207 J/AJ/124/1670 J/AJ/125/3302 J/AJ/126/1023 J/AJ/126/2421 J/AJ/126/3007 J/AJ/128/463 J/AJ/130/1719 J/AJ/131/1674 J/AJ/131/582 J/AJ/132/676 J/AJ/133/2258 J/AJ/133/439 J/ApJ/564/421 J/ApJ/607/426 J/ApJS/156/47 J/ApJS/94/749 J/PASP/118/659

Liechti S.: J/A+A/314/615 J/A+AS/102/611 Liefke C.: J/A+A/417/651 Lieu R.: II/203A IX/35 J/ApJS/93/569 Light R.M: II/246 Light R.M.: J/ApJ/448/179 J/ApJ/478/124 Lignieres F.: J/A+A/475/1053 Ligori S.: J/AJ/129/1534 Liivamaegi L.J.: J/A+A/459/L1 Liivamagi L.J.: J/A+A/462/811 Liller W.: J/A+A/459/875 J/A+AS/109/479 J/A+AS/109/487 LILLER W.: J/A+AS/112/439 Liller W.: J/A+AS/119/191 J/A+AS/125/1 LILLER W.: J/AJ/107/230

Liller W.: J/AJ/112/2004 J/AJ/112/2020 J/AJ/114/189 J/AJ/116/2415 J/AJ/99/1831 J/PAZh/21/596

J/PAZh/22/269 J/PAZh/22/764 Lilly S .J.: J/ApJ/517/148 Lilly S.: J/ApJ/499/112 J/MNRAS/350/785 Lilly S.J.: J/ApJ/411/501 J/ApJ/464/79 J/ApJ/481/49 J/ApJ/635/243 J/MNRAS/238/1171 VII/225A Lima Neto G.B.: J/A+A/377/428 J/A+A/407/31 Limboz F.: J/ApJ/569/720 Limon M.: J/ApJS/148/97 Limongi M.: J/A+A/390/961 J/AJ/118/2839 J/ApJ/608/405 Lin H.: J/AJ/121/1840 J/AJ/127/1883 J/ApJ/565/208 J/ApJS/129/475 J/ApJS/130/237 J/ApJS/158/161

VII/203 Lin W.: J/AJ/117/2757 J/AJ/126/1286 Lin Y.C.: J/ApJS/101/259 J/ApJS/107/227 J/ApJS/123/79 J/ApJS/94/551 Lin Y.-T.: J/ApJ/639/816 Lindahl J.: J/A+A/428/555 Lindblad P.O: J/A+A/295/585 Linde P.: J/A+A/465/357 LINDE P.: J/A+AS/110/533 Linde P.: J/A+AS/73/37 Lindegren L.: I/250 J/A+A/381/446 J/A+A/442/281 Linden-Vornle M.J.D.: J/AJ/127/3075 J/MNRAS/351/1290 Linder S.: J/MNRAS/352/478 J/MNRAS/357/819 Linders H.P.J.: J/A+AS/119/37 LINDGREN H.: III/105 Lindgren H.: J/A+AS/73/37 Lindgren M.: J/A+AS/115/295 Lindholm D.M.: J/ApJS/95/163 Lindler D.: J/ApJ/612/481 J/ApJ/633/871 Lindler D.J.: J/ApJS/146/209 Lindner U.: J/A+A/316/123 J/A+A/341/709 Lindqvist M.: J/A+A/391/967 J/A+AS/122/79 J/A+AS/124/509 J/A+AS/92/43 Lindroos K.P.: II/127 Lindroos O.: J/A+A/350/513 Lineweaver C.H.: J/ApJ/640/1051 Ling J.C.: J/ApJS/127/79 Ling J.F.: J/A+AS/143/335 J/AJ/119/3084 J/AJ/127/1181 Ling N.F.: J/ApJS/127/79 Linick J.P.: J/AJ/130/23 Link R.: J/AJ/125/1352 J/AJ/130/2677 Linke R.A.: VIII/10 VIII/28 LINNALUOTO S.: J/A+AS/103/183 Linnell A.P.: J/A+A/408/611 J/A+A/455/1037 Linnell Nemec A.F.: J/AJ/108/222 Linsky J.L.: J/A+A/310/173 J/A+A/415/331 J/AJ/116/941 J/AJ/121/337 J/ApJS/165/256 Linz H.: J/A+A/434/987 Lipari S.: J/AJ/110/1993 J/AJ/116/1 Lipkin Y.: J/A+A/459/557 Lipovetskij V.A.: VII/61A VII/61A

Lipovetsky V.A.: J/A+A/366/771 J/A+A/373/24 J/A+A/374/907 J/A+A/442/109 J/A+AS/131/295

J/A+AS/135/511 J/A+AS/137/299 J/A+AS/139/1 J/A+AS/142/417 J/AJ/121/66 J/AJ/123/1292 LIPOVETSKY V.A.: VII/172 VII/173 Lipovka A.A.: J/AZh/76/83 J/AZh/77/3 J/AZh/79/963 Lipovka N.M.: J/A+AS/101/447 J/A+AS/87/1 J/AZh/62/229 J/AZh/76/83 J/AZh/77/3 J/AZh/79/963

VIII/33 Lipscy S.J.: J/AJ/117/1708 Lipunova N.A.: II/205 Lique F.: J/A+A/451/1125 Lira P.: J/AJ/112/2408 J/AJ/130/1345 J/AJ/134/1103 J/ApJ/640/603 J/ApJS/162/1 Lis D.C.: J/ApJS/156/127 Liseau R.: J/A+AS/101/59 J/A+AS/136/471

Lisenfeld U.: J/A+A/402/151 J/A+A/436/443 J/A+A/449/937 J/A+A/462/507 J/A+A/470/505

J/A+A/472/121 Lisi F.: J/A+AS/120/521 J/A+AS/142/65 J/AJ/112/2786 Liske J.: J/AJ/130/1516 J/MNRAS/354/L31 VII/240 VII/240 Lisker T.: J/A+A/430/223 J/A+A/442/1023 J/ApJ/635/243 LISSANDRINI C.: J/A+AS/112/347 Lissandrini C.: J/A+AS/140/351 Lissauer J.J.: J/AJ/116/1998 J/ApJ/634/625 Lisse C.M.: J/ApJ/612/481 Lister M.L.: J/AJ/130/1389 J/AJ/130/1418 J/AJ/130/2473 J/AJ/131/1262 J/ApJ/427/125 J/ApJ/609/539

J/ApJS/155/33 J/PASJ/52/997 Lister T.A.: J/A+A/467/785 J/MNRAS/358/795 J/MNRAS/362/117 J/MNRAS/372/1117 J/MNRAS/379/647

J/MNRAS/379/816 Liszt H.S.: J/A+AS/52/63 Little A.G.: VIII/2 Littlefair S.: J/MNRAS/373/687 Littlefair S.P.: J/MNRAS/356/1583 J/MNRAS/358/341 J/MNRAS/375/1220 J/MNRAS/376/580 Little-Marenin I.R.: V/68A Liu C.: J/ApJS/162/1 LIU F.K.: J/A+A/278/6 Liu F.K.: J/A+A/381/757 LIU F.K.: J/A+AS/106/361 Liu F.K.: J/ApJ/614/91 Liu J.: J/A+AS/132/233 Liu J.-F.: J/ApJS/157/59 Liu J.R.: J/A+A/391/775 J/A+A/422/377 J/A+A/454/379 J/MNRAS/376/1707 Liu M.C.: J/ApJ/602/571 J/ApJ/611/270 Liu Q.: J/A+A/355/171 LIU Q.: J/A+AS/101/49 Liu Q.: J/A+AS/125/475 J/A+AS/136/139 Liu Q.Y.: J/A+AS/124/281 Liu Q.Z.: J/A+A/429/1125 J/A+A/442/1135 J/A+A/455/1165 J/A+A/469/807 V/106 Liu W.M.: J/AJ/126/1665 J/AJ/131/2538 Liu W.W.: J/A+AS/132/83 J/A+AS/136/455 J/ApJ/522/846 Liu X.-W.: J/A+A/404/545 J/A+A/442/249 J/MNRAS/368/1959 Liu Y.: J/A+A/466/63 Liu Y.-L.: J/other/PBeiO/18.7 Liu Z.L.: J/A+A/374/1056 Liubarsky I.: J/other/APh/26.282 Livanou E.: J/A+A/451/431 Livingston W.: J/PASP/107/1042 Livio M.: J/A+A/465/481 J/ApJ/607/665 J/ApJ/612/481 J/ApJ/613/200 L.Jetsu: J/A+A/383/197 Llorente A.: J/A+A/474/793 Lloyd C.: J/MNRAS/368/1757 Lloyd Evans T.: J/A+A/397/997 LLOYD L.A.: VII/109 Loaring N.: J/MNRAS/335/673 J/MNRAS/342/802

Loaring N.S.: II/268 J/A+A/476/1191 J/MNRAS/328/805 J/MNRAS/362/1371 J/MNRAS/369/156

J/MNRAS/378/995 VII/223 VII/241 Lobachev V.A.: IX/36 Lobanov A.P.: J/AJ/130/2473 Lobanov: J/A+A/364/391

Lobo C.: J/A+A/393/439 J/A+A/429/101 J/A+AS/111/265 J/A+AS/122/409 J/A+AS/128/67

J/A+AS/129/281 J/A+AS/139/525 J/A+AS/139/559 Lockman F.J.: J/ApJS/105/369 J/ApJS/140/331 J/ApJS/97/1 VIII/58 Loden K.: III/35 LODEN L.O.: III/33 Loden L.O.: III/35

Lodieu N.: J/A+A/395/813 J/A+A/425/927 J/A+A/436/853 J/A+A/440/1061 J/A+A/448/189

J/A+A/450/147 J/MNRAS/373/95 J/MNRAS/374/372 J/MNRAS/380/712 Loewe M.: J/ApJ/466/254 Lohmann P.: J/AN/321/81 Loiseau N.: J/A+AS/75/67 Loken C.: J/AJ/118/2014 Loktin A.V.: J/AZh/80/8 V/96 Lomb N.R.: I/86A Lombardi M.: J/A+A/462/L17 LONDON: I/33 I/35 Long D.: J/A+A/444/643 J/AJ/123/3409 J/AJ/124/1651 J/AJ/131/1674 J/AJ/132/676 J/PASJ/55/771 Long D.C.: J/AJ/121/2308 J/AJ/121/2331 J/AJ/123/567 Long G.M.: J/AJ/123/3409 Long K.S.: J/AJ/110/2739 J/ApJ/625/156 J/MNRAS/361/1091 Longair M.: J/MNRAS/344/887 Longair M.S.: J/MNRAS/204/151 Longhetti M.: J/A+A/345/419 J/A+A/433/497 J/A+AS/130/251 J/A+AS/136/519 LONGLEY D.P.T.: J/A+AS/93/557 Longmore S.N.: J/MNRAS/367/1609 LONGO G.: J/A+A/296/319 Longo G.: J/A+A/428/339 LONGO G.: J/A+AS/105/433 Longo G.: J/A+AS/109/341 J/A+AS/114/527 J/A+AS/117/39 J/A+AS/133/325 J/AJ/109/517 LONGO G.: VII/167 Lonsdale C.: J/A+A/474/473 J/AJ/129/1198 J/AJ/129/2074 J/MNRAS/371/1891 Lonsdale C.J.: J/A+A/467/565 J/A+A/467/73 J/PASP/113/10 LONSDALE C.J.: VII/109 Lopes D.F.: J/A+AS/133/299 Lopes P.A.A.: J/A+AS/136/531 J/AJ/125/1660 J/AJ/125/2064 J/AJ/128/1017 J/MNRAS/380/1608 Lopez A.: J/A+A/465/331 Lopez B.: J/A+AS/134/545 Lopez C.E.: I/277 I/288 J/AJ/131/1000 Lopez de Coca P.: J/A+A/328/235 J/A+AS/100/571 LOPEZ DE COCA P.: J/A+AS/101/421 J/A+AS/106/21 Lopez de Coca P.: J/A+AS/129/9 J/A+AS/144/469 LOPEZ DE COCA P.: J/A+AS/93/189 Lopez M.: J/A+A/390/39 J/A+A/406/L9 J/A+A/410/813 Lopez S.: J/A+A/458/427 J/ApJS/162/1 Lopez-Corredoira M.: J/A+A/391/519 J/A+A/399/457 J/ApJ/491/L31 Lopez-Cruz O.: J/ApJS/153/397 Lopez-Garcia Z.: J/A+A/367/859 J/A+AS/107/353 Lopez-Garica Z.: J/A+AS/137/227 Lopez-Gonzalez M.J.: J/A+A/328/235 J/A+A/359/597 J/A+A/407/1059 J/A+A/450/715 J/A+A/456/261

J/A+AS/144/469 J/MNRAS/353/310 Lopez-Marti B.: J/A+A/416/555 Lopez-Santiago J.: J/A+A/379/976 J/MNRAS/328/45 Lopez-Sepulcre A.: J/A+A/471/573 Lorens H.: J/other/AISAO/32.31 Lorente R.: J/A+A/390/1033 Lorenz E.: J/A+A/390/39 J/A+A/406/L9 J/A+A/410/813 Lorenzani A.: J/A+A/417/615 Lorenzetti D.: J/A+A/399/147 J/A+A/419/999 J/A+A/422/141 J/A+AS/101/59 J/A+AS/136/471 Lorenz-Martins S.: J/A+A/431/565 Lorimer D.: J/ApJ/615/897 Lorimer D.R.: J/ApJ/605/759 J/MNRAS/273/411 J/MNRAS/342/1299 J/MNRAS/352/1439 J/MNRAS/359/1524

J/MNRAS/360/974 J/MNRAS/372/777 Lotz J.M.: J/AJ/128/163 Loulergue M.: J/A+AS/102/57 J/A+AS/130/333

Loup C.: II/228A J/A+A/267/515 J/A+A/319/235 J/A+A/377/945 J/A+A/389/149 J/A+A/406/51

J/A+A/454/185 J/A+AS/125/419 J/A+AS/99/291 LOVAS F.J.: VIII/12 Loveday J.: J/AJ/121/2308 J/AJ/121/2331 J/AJ/122/503 J/AJ/123/567 J/AJ/126/2579 J/AJ/127/1811

J/ApJS/107/201 J/MNRAS/278/1025 J/MNRAS/312/557 J/MNRAS/372/425 J/MNRAS/375/68 J/PASJ/55/771 Lovell E.J.: J/PASJ/52/997 Lovell J.: VIII/34 Lovell J.E.J.: J/AJ/128/1570 J/ApJ/614/607 J/ApJ/616/110 J/ApJS/155/33 J/PASJ/55/351 Lovis C.: J/A+A/468/1115 J/A+A/474/293 Lovkaya M.N.: J/A+AS/139/555 Low F.J.: II/126 Lowe L.: VII/79 Lowe S.R.: J/A+A/474/1093 LOWENTHAL J.D.: J/ApJ/405/498 Lowenthal J.D.: J/ApJ/481/673 J/ApJ/505/50 Lowrance P.: J/AJ/126/3007 J/AJ/128/463 Lowrance P.J.: J/AJ/126/2421 Lowry S.C.: J/A+A/444/287 LOYOLA P.: I/43 Loyola P.: I/82 LOYOLA P.: I/89A Loyola P.: J/A+AS/100/489 LOYOLA P.: J/A+AS/105/429 Loyola P.: J/A+AS/109/501 J/A+AS/115/75 J/A+AS/126/311 LOYOLA P.: J/A+AS/95/355 Loyola P.R.: I/78 Lrzeminski W.: J/A+AS/120/139 Lu J.: J/A+AS/127/181 Lu L.: J/A+AS/118/1 J/A+AS/136/1 J/ApJ/472/509 J/ApJ/508/200 J/other/PBeiO/17.120 J/other/PBeiO/24.71 LU LIZHI: I/181 I/182 Lu L.Z.: J/A+A/430/327 Lu L.-Z.: J/other/PBeiO/17.105 Lu L.Z.: J/other/PBeiO/32.87 Lu N.Y.: J/ApJ/449/527 J/ApJ/473/822 J/ApJS/103/331 Lu P.: J/AJ/112/628 J/AJ/117/2757 Lu P.K.: J/AJ/126/762 J/ApJS/83/203 V/81 Lu Y.: J/ApJ/645/856 Lu Y.-J.: J/ApJ/645/856 Lub J.: J/A+AS/70/95 LUB J.: J/A+AS/85/915 Lub J.: J/A+AS/96/389 Lubin L.M.: J/AJ/116/560 J/AJ/116/584 J/AJ/120/1 J/AJ/122/1125 LUBOWICH D.A.: VI/56 Lucarelli F.: J/A+A/390/39 J/A+A/406/L9 J/A+A/410/813 Lucas P.W.: J/MNRAS/361/211 Lucas R.: J/ApJ/613/200 Lucas R.A.: J/AJ/112/1335 J/AJ/120/2747 J/AJ/127/131 J/ApJ/600/L155

Lucatello S.: J/A+A/377/123 J/A+A/404/187 J/A+A/420/683 J/A+A/440/901 J/A+A/454/581

J/A+A/455/271 J/A+A/464/927 J/A+A/464/967 J/AJ/128/2402 LucatelloS.: J/A+A/464/953 Lucey J.R.: J/AJ/128/1558 J/MNRAS/313/469 J/MNRAS/327/1004 J/MNRAS/327/249 J/MNRAS/327/265

J/MNRAS/369/1131 Luciani M.: J/A+AS/130/109 Luck E.R.: III/139A Luck R.E.: J/A+A/381/32 J/A+A/392/491 J/A+A/413/159 J/AJ/128/343 J/AJ/129/1063 J/AJ/129/433

J/AJ/130/1880 Ludendorff H.: IV/26 Ludwig H.-G.: J/A+A/442/281 Luedeke K.: J/A+A/336/920 Lueftinger T.: J/A+A/458/569 J/A+A/475/1053 Luetticke R.: J/A+AS/145/405 Lugaro M.: J/A+A/466/641 Lugger P. M.: J/AJ/115/708 Lugger P.M.: J/AJ/123/2261 J/ApJ/625/796 Luginbuhl C.B.: J/AJ/103/638 J/AJ/105/1571 J/AJ/127/3553 Luhman K.L.: J/AJ/131/1574 J/AJ/131/527 J/ApJ/525/466 J/ApJ/540/1016 J/ApJ/593/1093 J/ApJ/645/676

J/ApJS/151/387 Luinge W.: II/97 Lukac C.E.: I/90 LUKAC C.F.: I/73 Lumb D.: J/ApJ/554/742 Lumme K.: J/A+AS/113/115 Lumsden S.: J/ApJS/132/37 J/MNRAS/329/227 J/MNRAS/329/87 J/MNRAS/333/100 J/MNRAS/350/1485

J/MNRAS/352/939 Lumsden S.L.: J/A+A/461/11 J/A+A/474/891 J/MNRAS/274/1071 J/MNRAS/363/1111 J/MNRAS/366/1096

J/MNRAS/379/663 LUMSDEN S.L.: VII/175 Luna G.J.M.: J/A+A/435/1087 Lund C.: J/A+A/371/79 Lund N.: J/A+A/427/87 J/A+AS/129/1 J/A+AS/130/233 Lundbert J.: J/other/APh/26.282 LUNDGREN K.: J/A+AS/76/101 Lundgren S. C.: J/ApJS/121/171 Lundqvist P.: J/A+A/428/555 J/A+A/429/559 Lundstroem I.: J/A+A/433/185 J/A+A/465/357 Lundstrom I.: J/A+A/331/949 J/A+A/410/527 J/A+A/415/155 Lundstrom I. Stenholm B.: III/85 Lunine J.: J/PASP/118/1690 Luo C.-H.: J/A+A/474/783 Luo D.: J/other/PBeiO/16.18 Luo S.-G.: J/A+A/442/249 Luo S.G.: J/ApJS/159/282 Lupone S.: J/A+A/389/149 Lupton R.H.: J/A+A/423/755 J/AJ/121/2308 J/AJ/121/2331 J/AJ/121/31 J/AJ/122/503 J/AJ/123/3409

J/AJ/123/567 J/AJ/125/1817 J/AJ/126/1023 J/AJ/126/1720 J/AJ/126/2579 J/AJ/127/3034 J/AJ/127/3553 J/AJ/131/1 J/ApJS/155/257 J/ApJS/163/80

Luque-Escamilla P.: J/A+A/452/739 Luri X.: J/A+A/319/881 J/A+A/326/722 J/A+A/336/953 J/A+A/374/968 J/A+A/430/165 J/A+AS/140/55 Luthardt R.: J/A+A/323/113 Lutovinov A.: J/A+A/475/775 Lutovinov A.A.: J/PAZh/30/430 J/PAZh/30/589 Lutz D.: J/A+A/418/465 J/ApJ/566/880 J/ApJ/606/664 Lutz T.E.: J/AJ/108/222 LUTZ T.E.: V/7 Luu J.X.: J/AJ/131/527 Luyten W.: III/74 LUYTEN W.J.: I/130 I/130 Luyten W.J.: I/54A I/87B I/98A I/98A III/49 III/70 Ly C.: J/ApJ/648/7 J/PASJ/58/313 Lyke J.E.: J/AJ/127/3121 Lynas-GRay A.E.: J/A+A/282/899 Lynas-Gray A.E.: J/A+AS/129/505 J/MNRAS/380/1098 Lynch D.: J/A+A/459/875 Lynch D.K.: J/A+AS/147/5 J/AJ/121/362 Lynden-Bell D.: J/MNRAS/277/125 VII/176A Lynds B.: VII/9 Lynds B.T.: VII/7A Lynds R.: J/AJ/113/1691 J/AJ/116/146 J/ApJ/456/174 J/ApJ/459/L27 J/ApJ/468/633 J/ApJ/478/124 Lyne A.G.: J/MNRAS/273/411 J/MNRAS/301/235 J/MNRAS/328/17 J/MNRAS/328/855 J/MNRAS/335/275

J/MNRAS/342/1299 J/MNRAS/352/1439 J/MNRAS/353/1311 J/MNRAS/360/974 J/MNRAS/372/777 VII/156 VII/189 VII/8

LYNGA G.: III/5 III/6 III/7 J/A+AS/110/533 Lynga G.: J/A+AS/115/297 J/A+AS/73/37 LYNGA G.: VII/41 VII/66 Lyngaa G.: VII/92A Lyons R.W.: J/A+A/337/183 Lyra W.: J/A+A/453/101 Lytle D.: J/AJ/117/17 Lyubarsky Y.E.: J/ApJ/634/L89 Lyubchenko S.Yu.: J/AZh/79/328 Lyubchik Y.: J/A+A/416/655 Lyuty V.M.: J/A+A/376/775 J/AZh/79/1087 J/PAZh/21/652 J/PAZh/25/403 J/PAZh/26/13 J/PAZh/27/83

J/PAZh/27/942 Lyutyi V.M.: J/AZh/78/408 J/AZh/79/309 J/PAZh/23/113 J/PAZh/23/277 Lyytinen J.: J/A+A/383/197 M.: J/A+A/364/391 J/A+A/385/1014 J/A+A/390/707 J/A+A/404/775 J/A+A/421/937 J/A+A/440/487

J/A+A/448/983 J/A+A/466/1205 J/A+A/475/915 J/A+AS/120/139 J/A+AS/122/471 J/A+AS/128/459 J/MNRAS/273/1097 J/MNRAS/296/839

M. (PI); Barnstedt J.: VI/107 M. Zurita: J/A+AS/135/145 MA C.: I/151 Ma C.: I/251 J/A+AS/116/473 J/AJ/110/880 J/AJ/126/2562 J/AJ/127/3587 J/ApJS/141/13 Ma Huiyu: J/A+AS/114/337 J/A+AS/128/315 J/A+AS/129/569 Ma J.: J/A+A/397/361 J/A+A/449/143 J/A+AS/126/503 J/AJ/122/1718 J/AJ/126/1286 J/AJ/127/2579

J/AJ/129/1818 J/AJ/131/2245 J/PASP/118/1104 M.A.: J/A+A/356/529 VIII/73 Ma W.-Z.: I/277 I/288 Maas T.: J/A+A/463/243 MacAlpine G.M.: VII/171 VII/171 VII/171 Macau-Hercot D.: II/86 MACAU-HERCOT D.: III/39A Maccacaro T.: IX/13 MacCacaro T.: IX/15 Maccacaro T.: J/A+A/428/383 J/A+A/469/27 J/A+A/476/1191 J/ApJS/99/701 MacCacaro T.: VII/19 Maccagni D.: J/A+A/342/1 J/A+A/403/857 J/A+A/428/1043 J/A+A/439/413 J/A+A/457/79 J/A+A/467/73

J/A+A/474/473 J/A+AS/130/323 MACCAGNI D.: J/AJ/107/1629 Maccagni D.: J/MNRAS/382/279 VII/19 Maccarone M.C.: J/A+A/368/267 Maccarone T.J.: J/ApJ/586/814 J/ApJ/621/104 Macchetto D.: J/ApJS/107/215 J/ApJS/127/39 Macchetto F. D.: J/ApJS/122/81 Macchetto F.: J/A+A/287/769 J/A+A/288/413 Macchetto F.D.: J/A+AS/138/253 J/ApJ/513/733 J/ApJ/522/802 J/ApJ/551/973 MacConnel JACK D.: III/146 MacConnell D. J.: III/170 MacConnell D.J.: III/46 J/A+AS/133/201 Macconnell D.J.: J/AJ/77/486 MacConnell D.J.: J/AJ/89/1897 J/PASP/115/351 J/PASP/117/199 MacConnell J.D.: J/ApJS/86/453 Macdonald G.H.: J/A+A/325/1163 J/A+A/410/597 J/A+A/453/1003 J/A+AS/119/333 J/A+AS/133/29 Macdonald II W.: J/other/IBVS/5526 Mac-Donnell D.J.: III/15 III/76A MacDowall R. J.: J/A+AS/119/489 MACERONI C.: J/A+A/289/871 Maceroni C.: J/A+A/311/523 J/AJ/121/254 MacGillivray H.: J/A+A/342/1 MACGILLIVRAY H.: J/A+AS/112/347 MacGillivray H.: J/A+AS/117/519 J/AJ/121/2331 J/ApJS/140/239 MacGillivray H.T.: I/277 I/288 J/A+A/425/367 Macgillivray H.T.: J/MNRAS/362/689 Machado M.: J/A+A/477/877 Machado R.S.: J/AJ/126/1750 Machado-Palaez L.: J/PASP/111/335 Machalski J.: J/A+AS/128/153 J/ApJS/123/41 VIII/53 VIII/53 Maciejewski G.: II/264 II/264

Maciel W.J.: J/A+A/279/567 J/A+A/282/436 J/A+A/345/629 J/A+A/381/32 J/A+A/392/491 J/A+A/423/199 J/A+A/423/517 J/A+A/431/565 J/A+A/472/101 J/A+A/475/217 J/MNRAS/298/332

Mack J.: J/AJ/120/2747 J/ApJ/569/689 Mack K.-H.: J/A+A/406/579 J/A+A/435/863 J/A+AS/133/129 J/A+AS/139/359 J/MNRAS/329/877

J/MNRAS/348/857 J/MNRAS/370/1034 Mack P.: J/AJ/105/788 Mackay C.D.: J/A+A/287/769 J/A+A/288/413 Mackenty J.: J/AJ/117/206 MacKenty J.W.: J/AJ/118/1684 Mackey A.D.: J/MNRAS/345/747 MacLeod G.C.: J/A+AS/110/81 Macmillan D.S.: J/AJ/126/2562 J/AJ/127/3587 MacMillan D.S.: J/ApJS/141/13 Macomb D.J.: J/ApJS/120/335 MacRae D.A.: II/221A Macrae D.A.: J/ApJ/116/592 Macri L.: J/AJ/116/102 J/ApJ/523/540

Macri L.M.: J/AJ/121/861 J/AJ/121/870 J/AJ/126/175 J/AJ/131/527 J/ApJ/512/48 J/ApJ/514/614

J/ApJ/521/155 J/ApJ/529/723 Macrina S.: J/A+A/416/917 Madau P.: J/AJ/128/163 Madden S.C.: J/ApJ/599/1049 Maddox N.: J/AJ/126/63 Maddox S.: J/MNRAS/279/459 J/MNRAS/329/227 J/MNRAS/329/87 J/MNRAS/352/939 Maddox S.J.: J/ApJS/107/201 J/MNRAS/246/433 J/MNRAS/269/151 J/MNRAS/289/263 J/MNRAS/333/100

J/MNRAS/350/1485 J/MNRAS/362/1143 VII/221 Madej J.: J/A+A/407/631 J/A+A/420/507 J/A+A/430/1143

Madejsky R.: J/A+A/379/436 J/A+A/384/81 J/A+A/392/741 J/A+A/456/881 J/A+A/462/553

J/A+AS/110/131 Mader J.: J/ApJ/560/566 J/ApJS/129/435 Mader J.A.: J/AJ/127/3121 J/AJ/130/234 J/AJ/131/527 J/ApJ/640/1018 Mader S.: J/AJ/124/1954 J/AJ/128/16 J/AJ/129/220 J/MNRAS/362/689 Madgwick D.: J/MNRAS/329/227 J/MNRAS/329/87 J/MNRAS/352/939 Madore B.: B/bax Madore B. F.: J/ApJ/512/48 Madore B.F.: J/A+A/415/521 Madore B.F: J/AJ/106/66 Madore B.F.: J/AJ/106/964 J/AJ/128/2239 J/AJ/130/825 J/ApJ/463/26 J/ApJ/507/655 J/ApJ/508/491

J/ApJ/514/614 J/ApJ/521/155 J/ApJ/523/540 J/ApJ/529/723 J/ApJ/571/56 J/ApJ/619/L95 J/ApJ/633/871 J/ApJS/142/71 J/ApJS/147/29 J/ApJS/156/345 J/ApJS/164/38 J/ApJS/89/85

MADORE B.F.: J/ApJS/91/583 Madore B.F.: J/PASP/101/360 J/PASP/110/779 VII/170 Madsen G.J.: II/249 J/AJ/121/225 Madsen J.: J/other/APh/26.282 Madsen S.: J/A+A/381/446 Maeda A.: J/ApJS/162/428 Maeda Y.: J/ApJ/605/259 J/ApJ/635/214 Maeder A.: J/A+A/355/89 MAEDER A.: J/A+AS/101/415 J/A+AS/102/339 J/A+AS/103/97 Maeder A.: J/A+AS/115/339 J/A+AS/135/405 MAEDER A.: J/A+AS/96/269 MAEDER A. SCHALLER G.: J/A+AS/98/523 Maeder H.: J/A+A/428/1019 J/A+A/466/255 Maehara H.: II/106 II/106 MAEHARA H.: J/A+A/285/1 Maehara H.: J/other/PNAOJ/5.75 J/other/PNAOJ/6.1 J/other/PNAOJ/6.107 J/other/PNAOJ/7.37 Maercker M.: J/A+A/438/663 J/A+A/450/253 Maesano M.: J/A+A/327/61 J/A+A/395/11 J/A+A/402/151 J/A+A/423/935 J/A+A/451/435 J/A+AS/130/109 Magain P.: J/A+A/305/245 J/A+A/328/261 J/A+A/391/481 J/A+A/440/321 J/A+A/466/743 J/A+A/478/321 Magakian T.Y.: J/A+A/409/169 Magakian T.Yu.: J/A+A/399/141 J/A+A/463/1017 J/A+A/470/605 Magalhaes A.M.: J/ApJ/598/L43 J/ApJ/603/584 J/ApJS/141/469 Magazzu A.: J/A+AS/124/449 Magee D.: J/ApJS/150/1 Maggio A.: J/A+A/397/147 J/A+A/439/1149 J/A+A/459/199 J/ApJ/315/687 J/ApJ/348/253

Magliocchetti M.: J/MNRAS/333/100 J/MNRAS/341/993 J/MNRAS/350/1485 J/MNRAS/357/1267

J/MNRAS/379/260 Magnani L.: J/A+A/341/163 Magnenat P.: II/15 II/17 II/18 MagnenaT P.: II/38 Magnenat P.: II/39 II/7A Magneville C.: J/A+A/303/137 J/A+A/389/149 J/A+A/454/185 MAGNEVILLE C.: J/A+AS/109/447 Magnier E.: J/A+A/389/439 J/A+A/446/485 Magnier E.A.: II/208 J/A+A/278/36 J/A+A/352/228 J/A+A/373/63 J/A+AS/114/215 J/A+AS/126/401 Magnier E.A. Lewin W.H.G.: J/A+A/317/328 Magnitskii A.K.: J/PAZh/24/918 Magnussen N.: J/A+A/390/39 Magnusson P.: J/A+AS/113/115 Magorrian J.: J/AJ/129/2138 MAGRI C.: J/AJ/108/896 Magrini L.: J/A+A/367/498 J/A+A/379/90 Magris G.: J/AJ/127/1158 J/MNRAS/359/930 Magules L.: J/A+A/466/255 Mahabal A.: J/AJ/125/1660 J/AJ/125/2064 Mahabal A.A.: J/AJ/128/1017 Mahdavi A.: J/AJ/124/1266 J/AJ/130/1502 J/AJ/131/527 J/ApJ/518/69 J/ApJ/534/114 J/ApJ/554/L129

J/ApJ/607/202 J/ApJ/622/187 J/MNRAS/369/1375 Maheshwari V.U.: J/A+AS/117/557 Maheswar G.: J/A+A/402/963 J/MNRAS/355/1272 J/MNRAS/369/1822 J/MNRAS/379/1237 Mahmoudi W.F.: J/A+A/434/773 Mahoney J.: VIII/52B Mahoney S.: J/MNRAS/325/931 Mahoney T.J.: J/ApJ/491/L31 Mahoney W.: J/AJ/131/250 Mahoney W.A.: J/ApJS/127/79 Mahra H.S.: J/A+AS/122/111 Maia M.A.G.: J/AJ/113/1197 J/AJ/116/1 J/AJ/118/1131 J/AJ/123/1200 J/AJ/125/2307 J/AJ/126/1750

J/AJ/126/2268 J/AJ/128/2642 Maia M.R.G.: J/A+A/363/239 J/ApJ/578/943 Maihara T.: J/AJ/131/1608 J/PASJ/53/653 Maillard J.-P.: J/A+A/426/81

Mainieri V.: J/A+A/393/425 J/A+A/432/381 J/A+A/451/457 J/ApJS/139/369 J/ApJS/155/271

J/ApJS/155/73 J/ApJS/161/21 Maino D.: J/A+A/397/213 J/A+A/410/847 Maintz G.: J/A+A/442/229 J/A+A/442/381 J/A+AS/118/277 J/A+AS/144/227 Maintz M.: J/A+A/389/931 Mainzer A.K.: J/ApJ/597/555 Maio: J/A+A/421/937 Maio M.: J/A+A/430/603 Maiolino R.: J/A+A/409/115 J/A+A/409/79 J/A+A/418/465 J/A+A/434/801 J/A+A/459/85 J/A+A/466/31

J/A+A/468/807 J/ApJ/626/L15 J/ApJS/121/473 Maiorova E.K.: J/AZh/84/227 Maira A.: J/A+A/380/609 Maisala S.: J/A+A/383/197 Maisinger K.: J/MNRAS/341/1066 MAITRE V.: I/27

Maitzen H.M.: J/A+A/371/L5 J/A+A/373/153 J/A+A/385/867 J/A+A/395/823 J/A+A/403/937 J/A+A/412/721 J/A+A/441/631 J/A+A/443/157 J/A+A/444/941 J/A+A/454/171 J/A+A/454/179 J/A+A/458/293 J/A+A/459/871 J/A+A/462/591 J/A+AS/128/265 J/A+AS/128/573 J/A+AS/130/455 J/A+AS/133/1 J/MNRAS/362/1025

Maiz-Apellaniz J.: J/AJ/131/1184 J/ApJ/619/270 J/MNRAS/366/739 V/116 Majer P.: J/ApJ/315/687 Majewski S.R.: J/AJ/110/2747 J/AJ/120/284 J/AJ/121/935 J/AJ/125/1352 J/AJ/128/245 J/AJ/130/2677

J/ApJ/623/181 J/ApJS/109/45 Makarov D.I.: J/A+A/398/479 J/A+A/442/85 J/A+AS/139/97 J/AJ/127/2031 J/other/BSAO/50.5

J/PAZh/27/250 Makarov: I/275 Makarov V V: I/259 Makarov V.V.: I/250 I/276 IV/25 J/A+A/356/141 J/A+A/397/997 J/A+A/399/1167 J/A+AS/144/45

J/AJ/126/1996 J/AJ/129/2420 J/ApJ/595/1206 Makarova L.: J/A+AS/128/459 J/A+AS/133/181 J/A+AS/139/491 Makarova L.N.: J/A+A/396/473 J/A+A/433/751 J/AN/322/73 J/AN/325/307 MAKI T.: VII/140 Makidon R.: J/AJ/119/3026 J/AJ/127/2659 J/AJ/128/787 Makidon R.B.: J/AJ/116/1816 J/AJ/120/2747 J/AJ/123/1528 J/AJ/127/2228 Makinen K.: J/A+A/427/769 Makino F.: J/A+AS/121/461 Makishima K.: J/ApJ/518/656 J/ApJ/645/256 J/ApJS/133/1 J/ApJS/161/185 J/PASJ/56/965 Makovoz D.: J/AJ/131/250 Maksimov A.F.: J/A+A/385/87 J/A+A/422/627 J/A+AS/140/287 Maksym P.W.: J/ApJS/150/19 Malagnini M. L.: J/A+AS/126/267 Malagnini M.L.: J/ApJ/601/485 J/ApJ/613/312 J/ApJ/634/1319 J/PASP/113/1365 Malaguti G.: J/ApJ/607/L33 MalagutI G.: J/ApJS/106/399 Malanushenko V.P.: J/A+AS/105/125 Malaroda S.: III/204 III/216 III/249 J/A+AS/118/231 Malasan H.L.: J/PASJ/56/295 Malecki J.J.: J/AJ/128/1558 Malesani D.: J/A+A/474/793 Malhotra R.: J/PASP/118/1690 Malhotra S.: J/AJ/127/213

Malin D.F.: J/AJ/123/873 J/AJ/124/690 J/ApJ/274/534 J/MNRAS/346/787 J/MNRAS/355/1303

J/MNRAS/362/689 Malina R.F.: II/203A IX/35 IX/35 J/AJ/107/751 J/AJ/108/1843 J/AJ/130/825 J/ApJ/619/L95

J/ApJS/164/38 J/ApJS/93/569 Malizia A.: J/ApJ/607/L33 J/ApJ/636/765 J/ApJS/170/175 Malkan M. A.: J/ApJ/510/637 J/ApJS/125/349 Malkan M.: J/AJ/112/2408 J/ApJS/140/265 J/MNRAS/316/326 VII/157 MALKAN M.A.: J/A+A/298/375 Malkan M.A.: J/A+A/369/57 J/ApJ/453/616 J/ApJ/613/682 J/ApJ/616/707 J/ApJ/648/7 J/ApJS/142/79

J/ApJS/98/369 J/PASJ/58/313 Malkov O.: J/A+A/384/925 MALKOV O. Yu.: V/63 Malkov O.Yu.: II/171 J/A+A/320/79 J/A+A/353/958 J/A+A/446/785 J/BaltA/9/646 V/60 MALKOV O.Yu: V/75 Malkov Y. F.: J/ApJ/510/659 Malkov Y.F.: J/ApJS/115/185 Malkov Yu.F.: J/A+A/320/405 J/A+A/324/904 J/A+A/369/57 Mal’kov Yu.F.: J/AZh/74/853 J/PAZh/23/409 Mallama A.: J/A+AS/125/399 Mallen-Fullerton G.M.: J/ApJ/582/1123 Mallen-Ornelas G.: J/AJ/126/1208 J/ApJ/481/49 J/ApJ/582/1123 J/ApJ/627/621 Maller M.: J/MNRAS/367/1609 Mallik D.C.V.: J/A+AS/114/537 Mallik S.V.: J/A+A/409/251 J/A+AS/124/359 Mallmann F.: J/A+A/475/1125 Mallozzi R.S.: IX/20A J/ApJS/126/19 Malofeev V.M.: J/AZh/71/762 J/AZh/77/499 Maloney P.: J/ApJ/638/293 Malov I.F.: J/AZh/71/762 J/AZh/74/63 J/AZh/83/542 J/PAZh/29/571 Malov O.I.: J/AZh/74/63 J/AZh/77/499 J/AZh/83/542 M.altmann: J/A+A/447/173 MALUENDES S.: VI/59 Malumuth E.M.: J/AJ/104/495 J/AJ/107/1054 J/AJ/111/1128 J/ApJ/548/127 Malyuto V.: J/A+A/325/693 J/MNRAS/286/500 Mamajek E.: J/PASP/118/1690 Mamajek E.E.: J/AJ/124/1670 Mamon G.: J/A+A/363/62 VII/242 Mamon G.A.: J/A+A/387/1 J/A+A/466/481 J/A+AS/117/417 J/other/PASA/18.232 Mampaso A.: J/A+A/367/498 J/A+A/379/90 J/A+A/465/481 Manabe S.: J/PASJ/52/631 Manchado A.: J/A+A/458/173 J/A+A/458/533 J/A+AS/126/479 Manchester R.N.: J/AJ/129/1993 J/MNRAS/328/17 J/MNRAS/328/855 J/MNRAS/335/275 J/MNRAS/342/1299

J/MNRAS/352/1439 J/MNRAS/362/1189 J/MNRAS/372/777 VII/156 VII/189 VII/191 Mancini D.: J/A+A/428/1043 Mandushev G.I.: J/AJ/113/2104 MANFROID J.: II/170 II/188 II/200 Manfroid J.: II/202 J/A+A/364/689 J/A+A/463/981 J/A+AS/119/547 J/A+AS/123/353 J/A+AS/127/201 Mang O.: J/A+A/390/39 J/A+A/406/L9 J/A+A/410/813 Mangano V.: J/A+A/474/793 MANGENOT G.: I/63 Mangum J.G.: J/ApJS/81/5 Manimanis V.: J/A+A/374/1056 J/A+A/374/164 J/A+AS/133/7 Manimanis V.N.: J/A+A/369/960 MANITIUS K.: V/61 Mann G.: J/A+AS/119/489 Mann I.: J/A+AS/127/1 MANN P.J.: V/40 Mann R.G.: J/AJ/131/2383 J/MNRAS/343/293 J/MNRAS/344/887 J/MNRAS/351/1290 J/MNRAS/372/1621 Manners J.: J/AJ/129/2074 Manners J.C.: J/MNRAS/343/293 J/MNRAS/355/97 Mannheim K.: J/A+A/350/805 Mannings V.: J/ApJ/497/330 Mannone C.: J/A+AS/124/255 MANNONE R.: J/A+AS/110/177 Mannucci F.: J/A+A/434/801 J/A+A/468/807 Manoj P.: J/A+A/402/963 J/MNRAS/355/1272 Manrique W.T.: J/A+A/430/327 J/A+AS/118/1 J/A+AS/136/1 J/other/PBeiO/24.71 J/other/PBeiO/32.87 Manset N.: J/A+A/335/973 J/A+A/408/305 Manso Sainz R.: J/A+A/465/339 Mansoux B.: J/A+A/389/149 Mantegazza L.: J/A+A/363/29 J/A+A/367/759 J/A+A/385/793 J/A+A/445/901 J/A+AS/133/7 Manteiga M.: J/A+A/458/173 Mantovani F.: J/A+A/379/755 J/A+A/423/935 J/A+AS/120/89 J/A+AS/43/1 J/A+AS/82/41 Manzato P.: J/ApJ/569/720 Manzini F.: J/A+A/465/331 Mao P.H.: J/ApJ/596/944 J/ApJS/156/35 J/ApJS/165/19 Mao R.Q.: J/A+A/361/1079 Mao R.-Q.: J/A+A/461/197 Mao R.Q.: J/A+AS/127/181 J/other/PPMtO/14.185 Mao S.: J/ApJ/625/78 J/MNRAS/370/435 Maoz D.: J/AJ/121/3048 J/ApJ/409/28 J/ApJ/613/682 J/ApJ/632/799 J/ApJS/107/215 J/MNRAS/287/293

J/PASP/106/646 Maragoudaki F.: J/A+A/451/431 Maran S.P.: II/96 J/AJ/116/941 Marang F.: J/A+A/452/503 J/A+A/463/703 J/A+A/465/481 MARANG F.: J/A+AS/93/383 J/MNRAS/267/711 Marang F.: J/MNRAS/325/931 J/MNRAS/334/20 J/MNRAS/369/751 J/other/SAAOC/15.65 J/other/SAAOC/15.78 Marano B.: J/A+A/403/857 J/A+A/428/1043 J/A+A/457/79 Maraschi L.: J/A+A/327/61 J/A+A/439/413 J/A+A/467/73 Maraston C.: J/A+A/395/431 J/A+A/415/123 J/ApJ/621/673 J/MNRAS/369/697 Marcaide J.M.: J/A+A/370/409 MARCELIN M.: J/A+AS/113/35 Marcha M.J.: J/MNRAS/326/1455 J/MNRAS/329/877 Marcha M.J.M.: J/MNRAS/348/937 Marchal J.: VII/184 Marchenko S.G.: J/AZh/71/717 Marchenko S.V.: J/A+A/331/949 J/AJ/120/2101 Marchesini D.: J/AJ/134/1103 Marchev D.V.: J/A+A/424/993 Marcillac D.: J/A+A/423/867 J/A+A/451/57 Marciniak A.: J/A+A/465/331 Marco A.: J/A+A/372/477 J/A+AS/119/271 J/A+AS/123/1 Marcolino W.L.F.: J/A+A/477/877 Marconi A.: J/A+A/459/85

Marconi G.: J/A+A/310/715 J/A+A/330/453 J/A+A/380/478 J/A+A/414/503 J/A+A/425/509

J/A+A/462/L9 J/A+A/464/201 J/A+A/465/815 MARCONI G.: J/AJ/105/894 Marconi G.: J/AJ/109/173 J/AJ/118/1671 J/AJ/121/916 J/MNRAS/299/834 J/MNRAS/309/739 Marconi: J/A+A/469/529

Marconi M.: J/A+A/399/271 J/A+A/462/599 J/A+AS/121/327 J/AJ/129/2714 J/MNRAS/357/345

J/MNRAS/365/1357 J/MNRAS/371/1503 Marcout J.: V/100 V/100 V/127 MARCOUT J.: V/78 Marcout J.: V/84 MARCOUT J.: VI/14 Marcout J.: VII/24 Marcum P.M.: J/AJ/127/3213 Marcy G.W.: J/ApJ/634/625 J/ApJS/141/503 J/ApJS/152/261 J/MNRAS/370/163 J/MNRAS/372/163 Marechal L.J.L.: J/A+A/474/891 Marechal P.: J/A+A/328/617 MARECKI A.: III/121 Marecki A.: J/A+AS/135/273 Marengo M.: J/AJ/131/527 Marenzi A.M.: J/A+A/357/219 Marenzi A.R.: J/A+A/376/907 J/A+A/399/995 Margon B.: J/AJ/123/567 J/AJ/124/1651 J/AJ/126/1023 J/AJ/126/2209 J/AJ/126/2579 J/AJ/127/2838

J/AJ/129/2542 J/AJ/130/2439 J/ApJ/450/51 Margoniner V. E.: J/AJ/119/1562 Marie Jourdain de Muizon: III/242 Marigo P.: J/A+A/335/823 Marilli E.: J/A+A/333/205 J/A+A/388/298 J/A+A/454/301 J/AN/328/938 Marino A.: J/A+A/430/287 J/A+A/456/977 Marinoni C.: J/A+A/403/857 J/A+A/428/1043 J/A+A/457/79 J/ApJ/543/178 J/ApJ/625/6 Marinoni S.: J/A+A/456/467 MARKARIAN B.E.: VII/172 Markarian B.E.: VII/61A VII/61A VII/61A Markevitch M.: J/ApJ/504/27 Markkanen T.: J/A+A/372/276 Markov H.: J/A+AS/121/499 Markov V.N.: J/PAZh/33/143 MARKOVA N.: J/A+AS/108/561 Markova N.: J/A+AS/114/499 Markwardt C.B.: J/ApJ/633/L77 Marleau F.R.: J/AJ/131/250 J/ApJ/507/585 Marley M.S.: J/AJ/127/3553 Marlow D.R.: VIII/72 Maron O.: J/A+AS/147/195 Marouard M.: J/A+AS/137/451 Marquarding M.: J/AJ/123/873 J/AJ/124/1954 J/AJ/124/690 J/AJ/128/16 J/MNRAS/337/641 Marquette J.B.: II/279 Marquette J.-B.: J/A+A/377/945 Marquette J-B.: J/A+A/389/149 Marquette J.B.: J/A+A/454/185 Marquette J.-B.: J/MNRAS/371/1405

Marquez I.: J/A+A/285/404 J/A+A/366/771 J/A+A/370/409 J/A+A/374/907 J/A+A/393/389 J/A+A/416/475 J/A+A/420/873 J/A+A/429/101 J/A+A/442/109 J/A+AS/135/83 J/A+AS/142/417 J/MNRAS/351/1290

Marquez: J/A+A/445/939 Marr J.: J/PASP/97/932 Marraco H.G.: J/A+AS/144/195 MARRACO H.G.: J/AJ/105/258 MARRACO H.G: J/PASP/98/1133 Marrese P.M.: J/A+A/406/995 J/A+A/434/1107 Marrone D.P.: J/AJ/129/2856 MARS G.: VII/159 Marsakov V.A.: J/AZh/79/510 J/AZh/82/453 J/PAZh/30/173 MARSAKOV V.A.: V/57 Marsakov V.A.: V/58 MARSAKOV V.A.: V/89 Marsakova V.I.: J/A+A/366/100 Marschall L.A.: J/AJ/106/181 J/AJ/130/234 J/ApJ/640/1018 Marscher A.P.: J/A+A/476/L17 J/AJ/130/1418 Marsden B.G.: J/AJ/131/527 Marsden G.: J/MNRAS/372/1621 Marsden S.: J/A+A/475/1053 Marsh K.A.: II/246 J/ApJ/630/381 Marsh T.R.: J/A+A/466/1031 J/MNRAS/373/687 Marshall D.J.: J/A+A/453/635 MARSHALL E.E.: VII/181 Marshall F.E.: J/ApJ/253/485 J/ApJS/96/303 Marshall H.L.: J/AJ/107/751 Marshall J.: J/ApJ/640/768 Marshall J.L.: J/AJ/130/1916 Marshall S.: J/AJ/119/1901 J/AJ/127/2436 J/ApJ/542/257 J/ApJ/598/597 Marshall S.L.: J/AJ/114/326 J/AJ/127/334 J/ApJ/552/289 J/ApJ/562/337 Marsteller B.: J/A+A/439/129 J/AJ/130/2804 Marston A.P.: J/ApJS/111/163 Martayan C.: J/A+A/462/683 J/A+A/472/577 Martel A.R.: J/AJ/128/1990 J/ApJ/612/202 J/ApJ/621/53 J/ApJS/150/1 Martel F.: J/A+A/461/81 J/other/GCN/1070.1 Martel: J/ApJS/122/81 Martenis P.: IX/27 Marthinet M.C.: J/A+AS/117/467 Marti J.: J/A+A/452/739 J/A+A/472/557 Martic M.: J/A+A/478/197 Martin B.E.: J/A+AS/125/399 Martin C.: J/AJ/117/1890 Martin C.E.: J/MNRAS/353/1311 Martin C.L.: J/AJ/119/486 Martin D.C.: J/AJ/130/825 J/ApJ/619/L95 J/ApJ/633/871 J/ApJS/164/38 Martin E.: J/AJ/121/337 MARTIN E.L.: J/A+A/299/89 Martin E.L.: J/A+A/349/619 J/A+A/419/167 J/A+A/441/653 Martin E.L: J/A+A/446/485 MARTIN E.L.: J/A+AS/101/485

Martin E.L.: J/A+AS/123/473 J/A+AS/124/449 J/A+AS/128/485 J/AJ/126/1526 J/ApJ/546/1006

J/ApJ/602/571 Martin G.A.: VI/72 VI/72 Martin I.: J/A+A/376/1106 J/A+A/387/1146 Martin J. R.: J/PASP/111/812 Martin J.C.: I/224 I/225 I/226 Martin J.-M.: J/A+A/366/771 Martin J.-M: J/A+A/374/907

Martin J.-M.: J/A+A/419/469 J/A+A/430/373 J/A+A/442/109 J/A+A/465/71 J/A+AS/139/1

J/A+AS/142/417 Martin M.C.: J/A+AS/131/73 J/A+AS/131/77 MARTIN N.: III/105 III/111 III/147 Martin N.F.: J/MNRAS/375/831 J/MNRAS/380/281 Martin P.: J/A+A/268/419 J/A+A/321/363 J/A+A/358/819 J/A+A/413/159 J/ApJ/424/599 Martin P.G.: J/ApJ/644/1006 J/ApJS/126/85 Martin S.: J/A+A/336/920 J/A+A/372/588 J/A+A/455/971 Martin S.C.: J/AJ/105/1822 Martin V.A.F.: J/A+AS/137/269 Martin W.C.: VI/73 VI/74 VI/74 Martinez Fiorenzano A.F.: J/A+A/420/683 Martinez O.M.: J/A+A/422/925 Martinez P.: J/A+A/455/303 Martinez R.: J/A+A/462/621 Martinez-Gonzalez E.: J/MNRAS/311/456 MARTINEZ-ROGER C.: J/A+AS/107/365 Martinez-Roger C.: J/A+AS/117/227 J/A+AS/131/209 Martinez-Sansigre A.: J/MNRAS/372/1621 J/MNRAS/372/741 Martini P.: J/AJ/121/1840 J/ApJS/146/353 Martin-Luis F.: J/AJ/122/2055 J/AJ/125/3311 Martin-Mirones J.M.: J/MNRAS/311/456 Martin-Pintado J.: J/A+A/455/971 J/A+A/469/207 Martins D.H.: J/AJ/119/1793 Martins F.: J/A+A/457/637 J/ApJ/643/1011 Martins L.P.: J/MNRAS/357/945 Martmnez R.E.: J/A+A/355/138 Marvel K.: J/ApJS/107/175 Marvel K.B.: J/ApJS/120/147 Marx M.: J/A+AS/126/325 Mary D.L.: J/A+A/455/303 Marzari F.: J/A+A/420/683 Marziani P.: J/ApJS/145/199 Marzke R.O.: J/AJ/112/1803 J/AJ/122/2893 J/AJ/127/24 J/AJ/127/2455 J/ApJ/565/208 J/ApJS/111/163 Masana E.: J/A+A/426/827 J/A+A/450/735 J/A+AS/128/265 Masci F.: J/AJ/129/2074 Masci F.J.: J/AJ/131/250 J/PASP/113/10 Mase K.: J/other/APh/26.282

Masegosa J.: J/A+A/285/404 J/A+A/366/771 J/A+A/374/907 J/A+A/393/389 J/A+A/416/475 J/A+A/420/873 J/A+A/442/109 J/A+AS/137/419 J/A+AS/142/417 J/A+AS/91/285 J/MNRAS/351/1290 J/other/RMxAA/35.187

Maselli A.: J/A+A/451/435 Masetti N.: J/A+A/474/793 Masheder M.R.W.: J/A+A/315/578 J/A+A/357/1001 J/MNRAS/362/689 Mas-Hesse J.M.: J/AJ/110/212 Mas-hesse J.M.: J/ApJ/636/765 Mashonkina L.I.: J/AZh/81/209 Maslennikov K.L.: I/292 Maslowski J.: VIII/53 Masnou J.-L.: J/MNRAS/307/236 Mason B. D.: J/AJ/106/637

Mason B.D.: I/276 II/235 J/AJ/106/1639 J/AJ/106/352 J/AJ/108/2299 J/AJ/111/370 J/AJ/111/936 J/AJ/116/917 J/AJ/117/1890 J/AJ/118/1395 J/AJ/119/3071 J/AJ/119/3084 J/AJ/120/1120 J/AJ/121/3224 J/AJ/122/1586 J/AJ/123/1023 J/AJ/124/2254 J/AJ/127/539 J/AJ/128/3012 J/AJ/132/50 J/ApJ/525/420 J/PASP/107/299

Mason B.S.: J/MNRAS/354/485 Mason E.: J/A+A/459/875

Mason H.E.: J/A+A/370/1071 J/A+A/422/731 J/A+A/430/331 J/A+A/432/1137 J/A+A/433/731

J/A+AS/125/149 J/A+AS/131/153 Mason K.: J/MNRAS/362/1371 Mason K.O.: J/A+A/476/1191 J/MNRAS/274/1194 J/MNRAS/311/456 J/MNRAS/342/802 J/MNRAS/369/156 Mason: V/111 Masonkina L.I.: J/PAZh/28/41 MASSA D.: II/114 Massa D.: J/AJ/129/1642 Massaro E.: J/A+A/327/61 J/A+A/395/11 J/A+A/402/151 J/A+A/402/647 J/A+A/423/935 J/A+A/451/435

J/A+A/473/347 J/A+AS/124/123 J/A+AS/130/109 Massarotti A.: J/AJ/129/2294 Massarotti M.: J/A+A/389/787 J/A+A/425/797 Massey P.: II/236 III/116 Massey. P.: III/136

Massey P.: III/143 J/AJ/101/1408 J/AJ/103/1205 J/AJ/105/980 J/AJ/106/1906 J/AJ/108/1256 J/AJ/110/2715 J/AJ/113/1691 J/AJ/119/2214 J/AJ/121/2020 J/AJ/121/891 J/AJ/123/855 J/AJ/126/2867 J/AJ/131/2478 J/AJ/132/271 J/AJ/98/1305 J/ApJ/452/210 J/ApJ/454/151 J/ApJ/469/629 J/ApJ/490/328 J/ApJ/501/153 J/ApJ/505/793 J/ApJ/559/864 J/ApJ/576/880 J/ApJ/610/247 Massi F.: J/A+A/399/147 J/A+A/422/141 J/A+A/453/911 J/A+AS/136/471 MASSON C.R.: VIII/22 VIII/24 Massone G.: J/A+A/341/L79 J/A+AS/121/119 J/A+AS/130/359 MASSONE G.: J/A+AS/93/383 Masters K.L.: J/AJ/130/2613 Masuda K.: J/MNRAS/330/137 Masuda S.: J/PASJ/57/109 J/PASJ/57/13 Mateo M.: I/244A J/A+A/462/599 J/A+AS/112/407 J/A+AS/120/139 J/A+AS/122/471 J/A+AS/125/343

J/A+AS/128/19 MATEO M.: J/AJ/105/938 J/AJ/106/1508 Mateo M.: J/AJ/109/618 J/AJ/110/1141 J/AJ/110/1186 J/AJ/110/2166 J/AJ/115/1016 J/AJ/115/1840

J/AJ/115/1856 J/AJ/115/1894 J/AJ/115/592 J/AJ/117/2810 J/AJ/118/346 J/AJ/120/2496 J/AJ/123/279 J/AJ/123/3154 J/AJ/125/1261 J/AJ/125/2502 J/AJ/131/2114 J/AJ/131/912 J/AJ/132/347 J/ApJ/324/172 J/ApJ/562/303 J/ApJS/170/95 J/MNRAS/300/665 J/PASP/109/883 J/PASP/97/753

Mateos S.: J/A+A/444/79 J/A+A/469/27 J/A+A/474/473 J/A+A/476/1191 Matheson T.: J/AJ/130/2453 J/AJ/131/1648 J/AJ/131/527 J/ApJ/510/659 J/ApJ/594/1 J/ApJ/602/571

J/ApJ/613/200 Mathewson D.S.: II/34A J/ApJS/107/97 J/ApJS/81/413 Mathez G.: B/bax J/A+A/428/1043 J/A+A/457/79 MATHEZ G.: VII/88

Mathias P.: J/A+A/367/236 J/A+A/368/250 J/A+A/426/247 J/A+A/446/583 J/A+A/448/697

J/A+A/453/309 Mathieu R.: J/ApJ/566/993

Mathieu R.D.: J/A+A/339/423 J/A+A/382/899 J/A+A/409/1037 J/A+AS/130/245 J/AJ/117/2941

J/AJ/118/2409 J/AJ/121/2124 J/AJ/123/387 J/AJ/126/2922 J/PASP/106/281 Mathis J.S.: J/ApJ/635/560 Mathur S.: J/AJ/124/3042

Mathys G.: J/A+A/314/491 J/A+A/323/881 J/A+A/339/165 J/A+A/347/164 J/A+A/361/991

J/A+A/364/689 J/A+A/420/673 J/A+AS/123/353 Matilsky T.: IX/3 Matis H.S.: J/other/APh/26.282 Matkin N.V.: V/96 Matsubara Y.: J/MNRAS/330/137 Matsuda Y.: J/ApJ/648/7 J/PASJ/57/165 J/PASJ/58/313 Matsuhara H.: J/A+A/413/843 Matsumoto H.: J/PASJ/57/135 Matsumoto S.: J/MNRAS/321/77 J/PASJ/53/293 Matsunaga N.: J/MNRAS/353/705 J/MNRAS/364/117 J/MNRAS/370/1979 Matsuoka M.: J/other/GCN/1070.1 Matsushita S.: J/PASJ/57/135 Matsuura M.: J/A+A/376/997 J/MNRAS/376/313 Matsuyanagi I.: J/PASJ/58/L29 Matsuzaki K.: J/ApJS/134/77 Matt G.: J/A+A/409/79 J/A+A/466/31 J/A+AS/124/123 J/ApJ/570/100 J/ApJS/121/473 J/MNRAS/327/771 Mattei J.: J/A+A/323/113 Mattei J.A.: J/A+AS/124/143 J/AJ/130/776 J/other/IBVS/5526 Matteucci F.: J/A+A/352/117 J/A+A/445/93 J/A+A/459/875 J/MNRAS/336/879 Matthews H.: J/ApJ/639/259 Matthews H.E.: J/A+AS/92/43 Matthews J.: J/AJ/109/817 MATTHEWS J.M.: J/A+A/299/89 J/A+AS/101/485 Matthews J.M.: J/AJ/107/2093 MATTHEWS J.M.: J/AJ/107/679 Matthews J.M.: J/ApJ/459/278 Matthews K.: J/AJ/121/2189 J/AJ/121/97 J/AJ/123/1570 J/ApJ/564/421 Matthews L.D.: J/A+A/365/1 MATTILA K.: J/A+A/288/601 Mattila K.: J/A+A/405/999 MATTILA K.: J/A+AS/111/95 Mattila K.: J/A+AS/112/99 J/A+AS/119/153 J/A+AS/127/1 Mattila S.: J/A+A/429/559 Mattox J.R.: J/ApJ/481/95 J/ApJS/101/259 J/ApJS/107/227 J/ApJS/135/155 J/ApJS/94/551 Mattson C.J.: J/ApJS/103/433 Maturi M.: J/A+A/470/821 Matute I.: J/AJ/127/3075 J/MNRAS/351/1290 Mauas P.: J/A+A/326/249 Mauas P.J.D.: J/A+A/414/699 J/A+A/469/309 Maubon G.: J/A+A/366/68 Maubon: J/A+AS/136/509 Mauch T.: J/MNRAS/375/931 VIII/70A VIII/78 VIII/81 VIII/82 Maucherat: J/A+AS/31/439 MAURICE E.: III/105 Maurice E.: J/A+A/303/137 J/A+A/389/149 J/A+A/454/185 MAURICE E.: J/A+AS/109/447 Maurice E.: J/A+AS/30/261

Maurogordato S.: J/A+A/342/1 J/A+A/399/813 J/A+A/408/905 J/A+A/430/19 J/A+A/461/81

J/A+AS/130/323 J/ApJ/520/491 Mauron N.: J/A+A/314/585 J/other/Ap/46.475 Maury A.: J/A+A/413/1037 J/A+A/454/185 Mauskopf P.: J/ApJ/638/293 J/MNRAS/376/1073 Maximov A.F.: J/A+A/366/868 J/A+AS/105/503 J/AJ/127/1181 Maxted P.F.L.: J/MNRAS/356/89 J/MNRAS/363/529 Maxwell A.A.: J/AJ/132/393 May A.: J/ApJS/158/139 May J.: J/A+A/327/325 J/A+A/426/97 J/A+AS/115/81 VIII/39 MAYALL M.W.: III/182 MAYER C.J.: VIII/23 Mayer L.: J/ApJ/634/272 J/MNRAS/351/265 Mayer P.: J/A+A/403/675 J/A+A/463/1061 Mayer-Hasselwander H.A.: J/ApJS/101/259 J/ApJS/107/227 J/ApJS/123/79 J/ApJS/94/551 MAYER-HASSELWANDER H.A.: VII/120 Mayne N.J.: J/MNRAS/375/1220 J/MNRAS/376/580 Mayo Greenberg J.: III/130 Mayor: J/A+A/404/775 MAYOR M.: III/105 Mayor M.: J/A+A/282/831 MAYOR M.: J/A+A/299/53 Mayor M.: J/A+A/307/80 J/A+A/318/416 J/A+A/319/481 J/A+A/324/505 J/A+A/324/91 J/A+A/331/581

J/A+A/335/183 J/A+A/339/423 J/A+A/356/590 J/A+A/364/665 J/A+A/375/205 J/A+A/375/30 J/A+A/375/L27 J/A+A/379/999 J/A+A/388/632 J/A+A/390/267 J/A+A/392/215 J/A+A/395/97 J/A+A/398/363 J/A+A/406/373 J/A+A/410/1039 J/A+A/410/1051 J/A+A/414/351 J/A+A/415/1153 J/A+A/415/391 J/A+A/423/189 J/A+A/424/L31 J/A+A/426/695 J/A+A/427/313 J/A+A/430/165 J/A+A/438/251 J/A+A/449/723 J/A+A/458/873 J/A+A/458/997 J/A+A/466/743 J/A+A/469/783 J/A+A/471/L51 J/A+A/473/829 J/A+A/474/293 J/A+A/475/1125 J/A+A/476/1347

MAYOR M.: J/A+AS/108/25 J/A+AS/113/483 Mayor M.: J/A+AS/139/433 MAYOR M.: J/AJ/105/226 Mayor M.: J/AJ/114/1087 V/117 Mayor P.: J/A+A/320/74 Mayya Y.D.: J/AJ/108/1276 J/AJ/109/2503 Maza J.: J/A+A/416/475 MAZA J.: J/AJ/106/2575 Maza J.: J/AJ/112/2408 J/AJ/134/1103 J/ApJ/621/104 J/ApJS/162/1 Mazeh: J/A+A/404/775 Mazeh T.: J/A+A/375/L27 J/A+A/426/695 J/A+A/471/L51 J/A+A/475/1125 J/AJ/117/2941 J/AJ/124/1144

J/ApJ/466/415 J/MNRAS/370/1429 Mazets E.: J/ApJS/156/217 MAZODIER B.: I/74 Mazuk S.: J/A+A/459/875 J/A+AS/147/5 J/AJ/121/362 Mazumdar A.: J/A+A/463/243 Mazur B.: J/A+AS/118/303 J/AJ/110/2206 Mazure A.: J/A+A/258/243 J/A+A/310/31 J/A+A/310/8 J/A+A/381/420 J/A+A/428/1043 J/A+A/451/1159

J/A+A/457/79 J/A+A/459/679 J/A+A/461/81 J/A+AS/111/265 J/A+AS/114/565 J/A+AS/122/409 J/A+AS/129/399 J/AJ/120/1 J/ApJ/485/447

MAZURE A.: J/MNRAS/259/233 Mazure A.: J/MNRAS/382/279 Mazurier J.-M.: I/272 MAZURIER J.M.: I/63 Mazurier J.M.: J/A+AS/89/311 Mazzali P.A.: J/MNRAS/369/1880 Mazzarella J.M.: B/bax J/AJ/126/1607 J/AJ/127/3235 J/ApJS/62/751 J/PASP/110/779 Mazzei P.: J/A+A/363/517 J/A+A/454/453 J/A+AS/141/257 Mazzitelli G.: J/A+AS/133/403 Mazzotta P.: J/A+AS/133/403 Mc Call M.L.: J/A+A/336/667 McAdam W.B.: J/ApJS/80/137 J/MNRAS/268/602 VIII/1A McAlister H.A.: J/A+A/453/155 J/AJ/106/1639 J/AJ/106/352 J/AJ/106/637 J/AJ/108/2299 J/AJ/111/370

J/AJ/111/393 J/AJ/111/936 J/AJ/119/3084 McALISTER H.A.: VI/24 McAlister H.A.: VI/46 McArthur B.E.: J/A+A/476/73 McBreen B.: J/A+A/385/377 J/A+A/407/791 McBreen S.: J/A+A/385/377 McCabe C.: J/ApJ/636/932 McCall M.L.: J/AJ/116/2246 McCallon H.: II/156A II/246 McCammon D.: VI/62 McCAMMON D.: VII/43 Mccandliss S.R.: J/ApJ/636/1002 McCarthy C.: J/MNRAS/372/163 McCarthy D.D.: J/A+A/355/398 J/A+A/406/1135 J/AJ/110/880 McCarthy D.G.: V/64 McCarthy D.W.: J/AJ/121/1840 McCArthy D.W. Jr.: J/A+A/325/159 McCarthy D.W. Jr.: J/AJ/106/773 Mccarthy D.W.Jr: J/ApJ/604/141 Mccarthy D.W.JR: J/ApJ/630/206 McCarthy J.K.: J/A+A/336/263 McCarthy Jr D.W.: J/ApJS/94/749 McCarthy M.C.: J/A+A/459/317 McCARTHY M.F.: III/2 Mccarthy M.F.: J/AJ/89/636 McCARTHY M.F.S.J.: II/194A McCarthy P. J.: J/ApJS/122/81 McCarthy P.: J/A+A/431/793 Mccarthy P.J.: J/AJ/127/2455 McCarthy P.J.: J/ApJS/107/19 J/MNRAS/361/525 McCaughrean M.J.: J/A+A/392/239 J/A+A/395/813 J/A+A/436/853 Mccaughrean M.J.: J/A+A/440/1061

McCaughrean M.J.: J/A+A/448/189 J/A+A/450/147 J/A+A/458/461 J/ApJS/160/319 J/ApJS/160/353

J/ApJS/160/511 McCauley T.: J/other/APh/26.282 McClanahan T.: J/ApJS/156/217 McClintock J.E.: J/ApJ/381/33 J/ApJ/577/738 McClure R.D.: II/45 II/46 II/80 J/PASP/110/1172 J/PASP/116/1012 McClure-Griffiths N.M.: J/ApJ/626/887 McCoey C.: J/A+A/419/999 McCollough M.L.: IX/20A J/ApJS/144/213 J/ApJS/154/585 McConnell D.: J/MNRAS/289/225 McConnell M.: J/A+AS/143/145 McCook G.: J/A+A/463/1061 McCook G.P.: III/129 III/210 III/235A J/A+A/312/879 J/ApJ/439/983 McCracken H.J.: J/A+A/403/857 Mccracken H.J.: J/A+A/428/1043 McCracken H.J.: J/A+A/439/413 Mccracken H.J.: J/A+A/457/79 McCracken H.J.: J/MNRAS/350/785 McCulloch P.: J/PASJ/52/997 McCulloch P.M.: J/AJ/110/880 Mcculloch P.M.: J/AJ/127/3609 McCulloch P.M.: J/AJ/130/2529 McCumber M.P.: J/AJ/130/1083 McCuskey S.W.: II/221A Mccuskey S.W.: J/AJ/79/107 McDavid D.: II/220 J/A+A/304/235 J/A+A/383/580 McDermith R.J.: J/AJ/116/2549 McDonald J.S.: J/AJ/107/751 McDonald K.: J/AJ/107/751 J/AJ/108/1843 McDonald L.: J/ApJS/94/551 McDonald L.M.: J/ApJS/101/259 J/ApJS/107/227 J/ApJS/123/79 McDonald P.: J/ApJS/163/80 McDonald S.W.: J/AJ/103/983 J/AJ/104/862 J/AJ/109/1352 J/AJ/119/2008 J/AJ/119/936 Mcdowell J.: J/ApJS/161/271 McDowell J.C.: J/ApJS/95/1 McElroy D.B.: J/AJ/112/1335 J/ApJS/100/105 McELROY D.B.: V/44 McElroy D.B.: VII/72 Mcelwain M.W.: J/AJ/126/2487 J/AJ/127/2856 McEvoy J.: J/MNRAS/367/1609 Mcfadden M.T.: J/AJ/126/2048 McFadden M.T.: J/AJ/132/161 McGahee C.E.: J/AJ/132/161 McGale P.A.: J/MNRAS/274/1165 McGary R.S.: J/AJ/121/1192 MCGAUGH S.S.: J/AJ/106/530 McGee R.X.: J/other/PASAu/6.471 McGehee P.: J/AJ/131/1674 McGehee P.M.: J/AJ/123/3409 Mcgehee P.M.: J/AJ/128/426 McGilchrist M.M.: J/MNRAS/246/110 McGillivray H.: J/A+AS/130/323 McGILLIVRAY H.T.: II/130 McGillivray H.T.: J/A+AS/140/351 McGlynn T.A.: J/ApJ/616/1284 Mcgowan K.E.: J/AJ/127/2436 Mcgrath E.J.: J/ApJS/155/577 McGraw J.T.: J/AJ/112/742 J/ApJS/94/749 McGregor P.J.: J/AJ/126/2237 J/ApJS/69/99 McHardy I.: J/MNRAS/274/1194 Mchardy I.: J/MNRAS/362/1371 McHardy I.: J/MNRAS/369/156 VII/19 McHardy I.M.: J/MNRAS/342/802 Mchardy I.M.: J/MNRAS/352/131 McHardy I.M.: J/MNRAS/378/995 McIlroy P.: J/AJ/116/102 McIntosh D.H.: J/AJ/128/544 J/ApJ/619/193 McIntyre V.J.: J/MNRAS/329/227 McKay N.P.F.: J/MNRAS/328/17 J/MNRAS/335/275 McKay T.: J/AJ/119/1901 J/AJ/125/1817 McKay T.A.: J/AJ/119/2598 J/AJ/121/2308 Mckay T.A.: J/AJ/121/31 McKay T.A.: J/AJ/122/503 Mckay T.A.: J/AJ/127/1811 J/AJ/127/2436 McKay T.A.: J/AJ/130/968 J/AJ/131/621 J/AJ/131/633 Mckay T.A.: J/ApJ/598/260 McKay T.A.: J/ApJS/148/243 Mckean J.P.: J/MNRAS/377/430 McKean J.P.: VIII/72 McKEITH C.D.: J/A+A/293/703 McKenna F.C.: J/A+AS/139/135 Mckenzie E.: J/ApJ/639/816 Mckinstry K.: J/AJ/130/2584 McLACHLAN A.: II/130 McLaughlin B.M.: J/A+A/446/1185 McLaughlin D.E.: J/AJ/115/1801 Mclaughlin D.E.: J/AJ/127/105 McLaughlin D.E.: J/ApJ/559/812 J/ApJ/591/850 J/ApJS/161/304 McLaughlin M.A.: J/MNRAS/342/1299 Mclaughlin M.A.: J/MNRAS/352/1439 McLaughlin M.A.: J/MNRAS/359/1524 Mclaughlin M.A.: J/MNRAS/372/777 McLaughlin S.F.: V/62 McLEAN A.D.: VI/59 McLean B.: J/A+A/448/579 J/AJ/117/206 J/ApJ/559/606 J/ApJS/129/435 McLean B.J.: I/197A J/A+AS/130/359 McLean I.M.: J/ApJS/142/79 McLean I.S.: J/AJ/112/2500 J/ApJ/491/93 J/ApJ/581/258 Mclean I.S.: J/ApJ/597/555 Mcleod B.A.: J/AJ/130/1977 McLeod B.A.: J/AJ/131/527 Mcleod B.A.: J/ApJ/626/L5 McLeod B.A.: J/ApJS/96/117 Mclure R.J.: J/MNRAS/361/607 McLure R.J.: J/MNRAS/373/1531 McMahan R.: J/AJ/112/2408 McMahan R.K.: J/MNRAS/321/277 McMahan R.K.Jr: J/ApJS/106/1 J/MNRAS/292/499 J/MNRAS/305/259 McMahon R.: I/267 J/A+A/323/739 J/AJ/129/2074

McMahon R.G.: J/A+A/370/426 J/A+A/476/1191 J/AJ/112/407 J/AJ/121/1872 J/ApJS/126/133 J/ApJS/143/1 J/MNRAS/268/305 J/MNRAS/274/75 J/MNRAS/276/614 J/MNRAS/279/459 J/MNRAS/302/222 J/MNRAS/305/297 J/MNRAS/307/293 J/MNRAS/311/456 J/MNRAS/329/700

Mcmahon R.G.: J/MNRAS/343/293 J/MNRAS/346/1103 J/MNRAS/351/1290 J/MNRAS/358/333 McMahon R.G.: VII/221 McMaster M.: VII/149 MCMATH CHESTER M.: J/ApJS/93/145 McMillan R.: J/ApJ/416/62 Mcnally C.: J/AJ/127/180 Mcnamara B.: J/ApJS/161/9 McNamara B.J.: J/ApJ/595/187 Mcnamara B.J.: J/ApJ/602/264 McNamara B.R.: J/AJ/131/527 J/ApJ/502/558 J/ApJ/594/154 Mcnamara B.R.: J/ApJ/641/140 MCNAMARA D.H.: J/AJ/106/2493 McNeil R.C.: III/150A McNUTT D.P.: VII/80 McParland C.P.: J/other/APh/26.282 McSwain M.V.: J/ApJ/622/1052 Mcswain M.V.: J/ApJ/639/1069 McSwain M.V.: J/ApJS/161/118 McSweeney J.D.: IX/13 McVean J.R.: J/AJ/109/359 McWilliam A.: J/AJ/132/137 Meaburn J.: III/174 VII/232 MEAD J.M.: II/136 II/184 Mead J.M.: II/209 II/71 II/98 MEAD J.M.: IV/21 V/51 MEADE M.: II/114 Meade M.: J/ApJS/146/1 J/ApJS/146/125 Meade M.R.: II/67 J/AJ/118/1073 J/ApJ/635/560 J/ApJS/136/631 VI/105 Meadows S.A.: J/A+A/390/1023 Meakes M.: II/212A II/231 II/257 VII/149 Meatheringham S.J.: J/ApJS/75/407 J/ApJS/76/1085 Mebold U.: J/A+A/459/391 J/A+AS/126/325 VII/75 Medhi B.J.: J/A+A/469/713 Mediavilla E.: J/A+A/444/685 Medina Tanco G.A.: J/A+AS/136/261 J/A+AS/139/321 Medupe T.: J/A+A/472/1055 Medvedeva G.I.: II/140 II/205 Meech K.J: J/A+A/380/347 Meegan C.: J/ApJS/120/399 J/ApJS/156/217 Meegan C.A.: IX/20A J/ApJS/134/385 MEEKINS J.F.: VII/80 Meeus G.: J/ApJS/160/319 Megeath S.T.: J/A+A/466/989 J/AJ/126/1916 J/AJ/131/527 J/ApJ/638/897 Megeath T.: J/AJ/131/1574 Megier A.: J/ApJ/634/451 J/MNRAS/358/563 Mehlert D.: J/A+A/398/49 J/A+A/407/423 J/A+A/418/885 J/A+A/420/97 J/A+A/421/41 J/A+AS/141/449 Mehlert: J/A+A/395/753 Mehringer D.: J/ApJS/156/127 MEHTA S.: J/AJ/106/1533 Mei S.: J/AJ/131/143 J/ApJ/613/279 J/ApJ/644/30 J/ApJS/153/223 J/ApJS/154/509 Meibom S.: J/A+A/361/929 J/A+A/386/187 MEIDAV M.: VII/80 Meier D.L.: J/PASJ/52/997 Meijer J.: J/A+AS/93/463 Meijerink R.: J/A+A/404/L29 Meikle W.P.S.: J/MNRAS/369/1880 Meiksin A.: J/AJ/121/2308 J/AJ/121/2331 J/AJ/123/567 J/AJ/126/2579 J/MNRAS/372/425 J/MNRAS/377/787 Meinert D.: J/A+AS/120/77 Meisenheimer K.: II/253 J/A+A/285/785 J/A+A/377/442 J/A+A/396/513 J/A+AS/105/211 Meixner M.: J/ApJS/122/221 Mekarnia D.: J/A+A/466/L1 Mekkas M.: J/A+A/472/577 Melbourne J.: J/AJ/130/2584 Melbourne Obs.: I/148 Melchior A.L.: J/A+A/421/509 Melchior A.-L.: J/A+AS/145/11 MELCHIOR P.: I/47 Melekh B.Ya.: J/AZh/79/867 Melendez J.: J/A+A/475/1003 J/AJ/129/1607 J/AJ/129/303 J/AJ/132/137 J/ApJ/626/446 J/ApJS/124/527

VI/120 Melendez M.: J/A+A/410/359 J/A+A/469/1203 Melhuish S.J.: J/A+A/370/78 Meli A.: J/other/APh/26.282 Meliani M.T.: J/A+A/304/347 J/other/PASA/16.175

Mellier Y.: J/A+A/372/391 J/A+A/422/407 J/A+A/428/1043 J/A+A/439/413 J/A+A/444/365

J/A+A/457/79 J/A+AS/124/283 Mellilo J.F.: J/A+AS/125/399 Melnick J.: J/A+A/341/98 J/A+A/436/457 J/A+A/443/851 J/A+A/476/89 J/A+AS/125/247 J/A+AS/71/297

J/A+AS/91/285 J/AJ/110/463 J/MNRAS/366/480 Mel’nik A.M.: J/A+A/323/775 Melnik A.M.: J/PAZh/21/13 Melnikov S.: J/A+A/448/1075 Melnikov S.Y.: J/A+A/409/169 J/A+A/461/183 MELNIKOV S.Y.: J/A+AS/107/9 Mel’nikov S.Y.: J/PAZh/27/83 MEL’NIKOV S.YU.: J/A+A/282/L9 Mel’nikov S.Yu.: J/A+A/350/513 J/A+A/374/1049 MELNIKOV S.YU.: J/A+AS/103/183 Mel’nikov S.Yu.: J/A+AS/139/513 Mel’Nikov S.Yu.: J/AZh/78/408 Mel’nikov S.Yu.: J/PAZh/26/27 J/PAZh/27/942 Melnyk M.: II/156A Melo C.: J/A+A/356/590 J/A+A/375/130 J/A+A/458/895 J/A+A/458/997 J/A+A/461/509 J/A+A/466/743

J/MNRAS/367/1478 Melo C.H.F.: J/A+A/409/1037 J/A+A/460/695 Melo V.P.: J/ApJ/619/270 Melott A. L.: J/ApJS/122/355 Menanteau F.: J/AJ/128/1990 J/AJ/131/143 J/ApJ/612/202 J/ApJ/621/53 J/ApJ/644/30 J/ApJS/150/1 MENANTEAU F.: VII/205 Menard B.: J/A+A/410/33 Menard F.: J/A+A/349/619 Menci N.: J/A+A/392/395 J/A+A/437/883 Mendes de Oliveira C.: J/A+A/330/881 J/A+A/395/431 J/A+A/463/503 Mendes De Oliveira C.: J/AJ/130/55 Mendes de Oliveira C.: VII/213 VII/213 VII/213 Mendes L.T.S.: J/AJ/126/2971 Mendes S.O.: J/AJ/118/1131 Mendez M.E.: III/45 Mendez R.: J/A+AS/137/75 J/ApJS/162/1 Mendez R.A.: I/277 I/288 J/A+A/436/527 J/A+A/461/509 J/AJ/112/2408 J/AJ/126/762 Mendez R.H.: J/A+A/306/9 J/A+A/422/907 J/ApJ/635/290

Mendoza C.: J/A+A/410/359 J/A+A/414/377 J/A+A/418/1171 J/A+AS/123/159 J/A+AS/131/499

J/A+AS/135/159 J/A+AS/137/165 Menedez C.: J/A+AS/102/611 Meneux B.: J/A+A/403/857 J/A+A/428/1043 J/A+A/457/79 Mengel S.: J/A+A/443/41 Mennella A.: J/A+A/410/847 Mennessier C.: J/A+A/349/619 J/A+AS/129/337 Mennessier M. O.: J/A+AS/140/55 Mennessier M.O.: I/196 Mennessier M.-O.: J/A+A/317/761

Mennessier M.O.: J/A+A/319/881 J/A+A/326/722 J/A+A/336/953 J/A+A/374/968 J/A+A/390/173

J/A+AS/124/143 Mennickent R.E.: J/A+A/450/873 J/A+A/478/659 J/A+AS/121/113 J/MNRAS/350/679 J/MNRAS/361/1055 Men’shchikov A.: J/A+AS/126/39 Men’shchikov A.B.: J/A+A/318/879

Menten K.M.: J/A+A/393/115 J/A+A/418/103 J/A+A/428/1019 J/A+A/434/613 J/A+A/435/575 J/A+A/466/1065 J/A+A/466/989 J/A+A/475/925 J/A+A/476/1243 J/ApJ/606/664 J/ApJ/615/702 J/ApJS/166/526

Menzies J.: J/A+AS/126/297 MENZIES J.: J/MNRAS/267/711 Menzies J.W.: J/MNRAS/316/326 J/MNRAS/369/783 J/MNRAS/370/1979 J/MNRAS/376/313 Meola G.: J/A+A/372/862 Mera D.: J/A+A/328/83 Merand A.: J/A+A/433/1155 J/A+A/453/155 J/A+A/453/309 J/A+A/476/73 Mercer E.P.: J/ApJ/635/560 Merck M.: J/ApJS/101/259 J/ApJS/107/227 Mercurio A.: J/A+A/397/431 J/A+A/424/79 J/A+A/425/797 J/A+A/428/339 J/MNRAS/358/1116 Mereghetti S.: J/A+A/368/835 J/A+A/458/21 Merelli A.: J/AJ/123/567 Mergen J.: J/AJ/107/2120

Merighi R.: J/A+A/342/1 J/A+A/428/1043 J/A+A/457/79 J/A+AS/123/219 J/A+AS/130/323

J/AJ/110/2319 Merin B.: J/A+A/365/110 J/A+A/378/116 J/A+A/379/564 J/ApJ/638/897 Merkl: J/A+A/395/753 Merkulova N.I.: J/A+A/371/79 J/A+A/437/87 J/ApJ/622/129 J/other/Ap/48.304 J/PAZh/27/942

Merluzzi P.: J/A+A/389/787 J/A+A/397/431 J/A+A/424/79 J/A+A/425/797 J/A+A/428/1043

J/A+A/457/79 J/MNRAS/358/1116 Mermilliod J.C.: I/196 MERMILLIOD J.C.: II/104 II/105 II/105 Mermilliod J.C.: II/122B MERMILLIOD J.C.: II/124A II/163 Mermilliod J.C.: II/164 MERMILLIOD J.C.: II/165 II/166 II/167 Mermilliod J.C.: II/168 Mermilliod J.-C.: II/193 Mermilliod J.C.: II/29 II/35 II/63A II/89 MERMILLIOD J.C.: III/55 III/97A IV/1 Mermilliod J.C.: IV/17A MERMILLIOD J.C.: IV/2 IV/7 MERMILLIOD J.-C.: J/A+A/299/53 Mermilliod J.-C.: J/A+A/307/80 J/A+A/319/481 J/A+A/320/74 J/A+A/324/91 J/A+A/329/101 Mermilliod J.C.: J/A+A/331/81

Mermilliod J.-C.: J/A+A/333/897 J/A+A/335/183 J/A+A/339/423 J/A+A/345/471 J/A+A/354/881 J/A+A/360/509 J/A+A/361/945 J/A+A/375/30 J/A+A/375/989 J/A+A/399/105 J/A+A/399/543 J/A+A/423/189 J/A+A/461/509 J/A+A/473/829 J/A+AS/107/39 J/A+AS/134/301 J/AJ/107/2093 J/MNRAS/377/1737

MERMILLIOD J.C.: VI/5 MERMILLIOD M.: II/107 II/128 II/158A II/167 Mermilliod M.: II/215 II/57 II/63A II/74 MERMILLIOD M.: III/165 Merrelli A.: J/AJ/121/2308 J/AJ/121/2331 Merrelli A.J.: J/ApJS/126/209 Merrett H.R.: J/MNRAS/369/120 J/MNRAS/369/97 Merrifield M.R.: J/MNRAS/369/120 J/MNRAS/369/97 Merrill K.M.: J/AJ/106/1906 J/AJ/107/2120 J/AJ/116/1816 Merritt D.: J/ApJ/613/279 J/ApJ/613/682 J/ApJS/153/223 J/ApJS/154/509 Messarius T.: J/other/APh/26.282

Messina S.: J/A+A/364/205 J/A+A/367/910 J/A+A/393/225 J/A+A/400/659 J/A+A/409/1017

J/A+A/410/671 J/A+A/469/713 J/A+AS/138/87 Messineo M.: II/243 J/A+A/393/115 J/A+A/418/103 J/A+A/435/575 J/A+A/444/831 Meszaros P.: J/other/APh/26.282 Meszaros Sz.: J/A+A/394/943 J/A+A/465/943 J/other/P+SS/53.925 Metanomski A.: J/A+A/326/211 Metanomski A.D.F.: J/A+AS/131/197 Metcalfe L.: J/A+A/357/839 J/A+A/365/545 J/A+A/407/791 J/AJ/112/241 Metcalfe N.: J/MNRAS/236/207 J/MNRAS/267/431 VII/42A Metcalfe T.S.: J/AJ/131/582 Metchev S.: J/PASP/118/1690 Metevier A.J.: J/ApJS/126/209 Metlov V.G.: J/PAZh/21/652 Metlova N.V.: J/AZh/78/408 J/AZh/79/309 J/PAZh/28/764 Metreveli M.D.: III/230 Metrevelii M.D.: III/253 Metz M.: J/A+A/413/771 Metzger M.R.: J/AJ/115/635 J/AJ/116/469 Meurer G.: J/ApJ/606/L25 J/ApJ/621/53 J/ApJS/150/1 Meurer G.R.: J/AJ/128/1990 J/ApJ/612/202 J/ApJ/644/30 J/ApJS/165/307 Meurs E.: J/A+A/469/503 J/A+A/474/793

Meusinger H.: J/A+A/332/575 J/A+A/392/851 J/A+A/442/211 J/A+AS/137/305 J/A+AS/139/141

J/AJ/127/156 Mewe R.: J/A+A/376/1113 J/A+AS/97/443 Meyer B.S.: J/ApJ/611/452 Meyer C.: J/A+AS/110/107 J/A+AS/125/399 Meyer D.M.: J/ApJ/480/568 Meyer H.: J/A+A/390/39 J/A+A/406/L9 J/A+A/410/813 Meyer M.: J/AJ/124/1954 J/AJ/128/16 J/ApJS/165/307 Meyer M.J.: J/ApJ/633/871

Meyer M.R.: III/240 J/AJ/114/198 J/AJ/116/1816 J/AJ/124/1001 J/AJ/124/1670 J/AJ/126/1665

J/AJ/127/1131 J/AJ/130/1733 MEYER M.R.: J/ApJ/397/520 Meyer M.R.: J/PASP/118/1690 Meyer R.D.: J/AJ/117/548 J/AJ/127/1727 Meyer S.S.: J/ApJS/148/97 Meyer: VIII/73

Meylan G.: J/A+A/287/769 J/A+A/324/505 J/A+A/391/945 J/A+A/419/533 J/A+A/464/L21

J/AJ/114/1087 J/ApJ/409/28 Meynadier F.: J/A+A/381/941 J/A+A/422/129 J/A+A/436/117 Meynet G.: J/A+A/337/779 J/A+A/448/955 MEYNET G.: J/A+AS/101/415 J/A+AS/102/339 J/A+AS/103/97 Meynet G.: J/A+AS/115/339 J/A+AS/128/471 J/A+AS/135/405 MEYNET G.: J/A+AS/96/269 J/A+AS/98/523 Meyssonnier N.: III/190B J/A+AS/102/251 J/A+AS/102/451 Mezzetti M.: J/ApJ/569/720 Micela G.: J/A+A/341/751 J/A+A/345/521 J/A+A/355/651 J/A+A/397/147 J/A+A/399/919 J/A+A/404/927

J/A+A/412/495 J/A+A/416/281 J/A+A/417/945 J/A+A/421/175 J/A+A/423/1029 J/A+A/423/677 J/A+A/426/1007 J/A+A/430/287 J/A+A/430/941 J/A+A/445/673 J/A+A/450/993 J/A+A/455/903 J/A+A/456/977 J/A+A/459/199 J/A+A/460/133 J/A+A/461/977 J/A+A/462/123 J/A+A/462/245 J/A+A/463/275 J/A+A/464/211 J/A+A/468/353 J/A+A/474/495 J/ApJ/325/798 J/ApJ/348/557 J/ApJ/351/492 J/ApJ/414/846 J/ApJ/582/382 J/ApJ/588/1009 J/ApJ/606/466 J/ApJ/608/781 J/ApJS/102/75 J/ApJS/160/319 J/ApJS/160/353 J/ApJS/160/450 J/MNRAS/326/1161

Miceli A.: J/AJ/130/2453 J/AJ/131/1648 Michael C.D.: J/ApJS/99/223 Michael R.R.: J/AJ/114/1471 Michalowski T.: J/A+A/465/331 Michalska G.: J/A+A/434/89 Michalski D.E.: VI/105 Michard R.: J/A+A/383/30 J/A+A/423/833 J/A+A/429/819 J/A+A/441/451 J/A+AS/116/429 VII/184 Michaud K.: J/ApJS/115/1 MICHEL A.: VII/144 Michel A.: VII/164 VII/193 Michel E.: J/A+A/346/586 J/A+A/468/637 J/A+A/478/197 Michel R.: J/A+A/275/201 J/A+A/422/527

Michelson P.F.: J/ApJ/590/109 J/ApJ/609/564 J/ApJ/626/95 J/ApJS/101/259 J/ApJS/107/227

J/ApJS/123/79 J/ApJS/94/551 Mickaelian A.M.: III/237A J/A+A/414/487 J/A+A/426/367 J/A+A/449/425 Mickaelian A.M: J/other/Ap/42.1 Mickaelian A.M.: VI/116 Micol A.: J/A+AS/109/267 Middour J.: J/AJ/112/2319 Mieske S.: J/A+A/383/823 J/A+A/418/445 J/A+A/463/503 J/A+A/472/111 Mighell K.J.: J/AJ/110/1649 J/AJ/111/1596 J/AJ/114/1458 J/AJ/116/2395 J/AJ/124/601 J/ApJ/462/241

J/ApJ/617/L41 Migliori L.: J/A+AS/94/327 Mignani R.: J/A+A/384/81 Mignani R.P.: IX/34 J/A+A/379/436 J/A+A/379/740 J/A+A/392/741 J/A+A/408/1029 J/A+A/458/245 Mignano A.: J/A+A/456/881 J/A+A/457/517 J/A+A/462/553 Mignard F.: J/A+A/459/669 J/A+A/462/795 J/A+AS/135/231 Migniani R.P.: J/A+A/470/317 Mignoli M.: J/A+A/342/1 J/A+A/392/395 J/A+A/409/79 J/A+A/432/69 J/A+A/437/883 J/A+A/457/501

J/A+A/466/31 J/A+A/466/41 J/A+AS/130/323 J/AJ/127/3075 J/ApJ/510/659 MIHAJLOV A.A.: J/A+AS/103/57 MIHALAS D.: VI/80 Mihos J.C.: J/ApJ/582/170 Mikayelyan G.M.: VI/116 Mikhail A.: J/AJ/127/1131 Mikhailov V.: J/ApJS/115/185 Mikhailov V.P.: J/A+A/376/775 J/PAZh/27/83 J/PAZh/27/942 Mikles V.J.: J/ApJS/155/73 Miknaitis G.: J/AJ/130/2453 J/AJ/131/1648 J/ApJ/602/571 Mikolajewska J.: J/A+A/327/191 J/A+A/463/703 J/A+AS/146/407 Mikolajewski M.: J/A+A/415/273 J/A+AS/116/1 Milano L.: J/A+A/318/797

Miley G.: J/A+A/362/519 J/A+AS/143/303 J/AJ/123/637 J/MNRAS/302/222 J/MNRAS/305/297

J/MNRAS/351/1290 Miley G. K.: J/ApJS/122/81 Miley G.K.: J/A+A/402/171 J/A+A/406/593 J/A+A/431/793 J/A+A/461/823 MILEY G.K.: J/A+AS/108/61

Miley G.K.: J/A+AS/108/79 J/A+AS/124/259 J/A+AS/70/95 J/A+AS/96/389 J/AJ/128/1990

J/ApJ/612/202 J/ApJ/621/53 J/ApJS/150/1 J/ApJS/158/139 Millan-Gabet R.: J/ApJ/620/961

Millar T.J.: J/A+A/325/1163 J/A+A/358/257 J/A+A/453/1003 J/A+AS/119/333 J/A+AS/121/139

J/A+AS/133/29 J/ApJS/117/427 Miller A.C.: V/102 V/105 V/109 V/95 Miller B.W.: J/AJ/114/2381 J/AJ/127/2674 Miller C.: J/ApJS/115/1 J/ApJS/148/243 J/MNRAS/360/587 Miller C.J.: J/AJ/124/1918 J/AJ/130/968 J/AJ/131/1288 J/AJ/131/2018 J/ApJ/520/491 J/PASJ/55/771 Miller E.W.: J/PASP/86/829 Miller G.J.: J/AJ/122/248 Miller H.R.: J/ApJS/115/185 Miller J.: J/ApJS/81/5 J/ApJS/83/203 Miller K.: J/PASJ/52/997 Miller L.: II/211 II/268 MILLER L.: J/A+AS/112/347 Miller L.: J/A+AS/140/351 J/ApJ/505/L7 J/ApJ/511/612 J/MNRAS/328/805 J/MNRAS/335/673 VII/223

VII/241 Miller M.: J/A+A/461/197 Miller N.A.: J/AJ/123/3018 J/AJ/124/2453 J/AJ/125/2393 J/AJ/125/2427 J/AJ/130/2541 J/AJ/130/47

J/AJ/131/2426 J/ApJ/613/841 J/ApJS/134/355 MILLER P.: J/MNRAS/263/425 Miller P.: J/MNRAS/307/293 Miller R.: J/other/Obs/116.382 Miller S.: J/A+AS/135/359 Milliard B.: J/A+A/303/661 J/AJ/130/825 J/ApJ/619/L95 J/ApJS/164/38 J/MNRAS/300/303 J/MNRAS/312/442 Millington S.J.C.: J/MNRAS/240/785 Millis R.L.: J/AJ/109/1363 Mills B.Y.: VIII/2 Milne M.: J/AJ/128/712 Milne P.: J/A+A/428/555 Milone A.A.E.: J/AJ/108/1987 J/AJ/109/269 Milone E.F.: J/A+A/452/959 J/AJ/109/359 J/AJ/110/2400 J/PASP/106/1120 Milosavljevic M.: J/ApJ/613/279 J/ApJS/153/223 J/ApJS/154/509 Milovanovic N.: J/A+A/365/656 J/A+A/417/375 J/A+A/441/391 Milsztajn A.: J/A+A/303/137 J/A+A/389/149 J/A+A/454/185 MILSZTAJN A.: J/A+AS/109/447 Milvang-Jense B.: J/A+A/444/365 Milvang-Jensen B.: J/A+A/427/397 Milvang-jensen B.: J/ApJ/630/206 Milvang-Jensen B.: J/MNRAS/361/109 J/MNRAS/366/144 J/MNRAS/371/1777 Mimica P.: J/A+A/454/855 Minakov A.A.: J/A+A/420/447 Minchin R.F.: J/AJ/123/873 J/AJ/124/1954 J/AJ/124/690 J/AJ/128/16 J/MNRAS/337/641 J/MNRAS/342/738

J/MNRAS/346/787 J/MNRAS/355/1303 MINDEZ R.: J/AJ/106/2575 Mine Y.: J/PASJ/51/791 Mineo T.: J/A+A/368/267 J/A+A/448/L9 J/A+A/474/793 Minezaki T.: J/ApJS/141/23 J/ApJS/141/31 Mingaliev M.: J/AZh/72/12

Mingaliev M.G.: J/A+A/370/78 J/A+A/469/451 J/A+AS/139/545 J/AZh/76/83 J/AZh/77/407

J/AZh/80/978 J/AZh/82/420 J/AZh/83/241 J/AZh/84/227 J/AZh/84/387 J/other/BSAO/46.28 Minier V.: J/A+A/403/1095 J/A+A/432/737 J/MNRAS/363/405 J/MNRAS/367/1609 J/MNRAS/368/1223 Minikhulov N.Kh.: J/A+A/327/145 Minikulov N.Kh.: J/AZh/77/420 Mink D.J.: J/PASP/111/438 Minnit D.: J/A+AS/107/39 Minniti D.: J/A+A/303/468 J/A+A/406/75 J/A+A/423/97 J/A+A/435/657 J/A+A/465/799 J/A+A/466/157

J/A+AS/113/299 J/AJ/114/326 J/AJ/127/2114 J/AJ/127/24 J/AJ/127/334 J/AJ/128/1167 J/ApJ/542/257 J/ApJ/552/289 J/ApJ/562/337 J/ApJ/582/1123 J/ApJ/598/597 J/ApJ/600/716 J/ApJ/640/L43 J/MNRAS/343/665

Minniti M.: J/A+A/474/121 Minns A.R.: J/MNRAS/315/839 Minor R.H.: J/other/APh/26.282 Minter A.: J/AJ/119/2801 J/PASJ/52/997 Miocinovic P.: J/other/APh/26.282 Mirabel I.F: J/A+AS/129/455 Mirabel I.F.: J/AJ/110/1993 Mirabel I.R.: J/A+A/429/1125 Miralles J.-M.: J/A+A/452/119 J/A+A/456/881 J/A+A/462/553 Miranda L.F.: J/A+A/365/110 J/A+A/378/116 J/A+A/379/564 J/A+A/462/591 Mirioni L.: J/A+A/476/1191 Mironov A.V.: J/AZh/72/80 Miroshnichenko A.S.: J/A+A/397/1035 J/A+A/408/305 J/A+AS/131/469 J/A+AS/147/5 Mirtadjieva K.: J/A+A/440/53 Mirzoyan R.: J/A+A/390/39 J/A+A/406/L9 J/A+A/410/813 Misanovic Z.: J/A+A/426/11 J/A+A/448/1247 J/A+A/451/835

Mishenina T.V.: J/A+A/418/551 J/AZh/80/458 J/AZh/84/429 J/AZh/86/158 J/MNRAS/371/879

J/PAZh/26/465 Mishurov Yu.N.: J/A+A/323/775 Miskotte K.: J/A+AS/128/179 Missana M.: J/AJ/109/1903 Missana N.: J/AJ/109/1903 Misselt K.: J/ApJS/154/155 Misselt K.A.: J/ApJS/143/377 Missoulis V.: J/MNRAS/351/1290 Miszalski B.: V/127 Mitchell D.: J/MNRAS/358/795 Mitchell D.A.: J/AJ/130/1358 Mitchell E.: J/MNRAS/344/283 Mitchell G.F.: J/ApJ/559/307 J/ApJ/639/259 Mitchell K.J.: J/AJ/87/1429 J/ApJS/153/119 MITCHELL K.J.: VII/63 VIII/26 Mitchell R.I.: II/52 II/5A II/84 Mito H.: J/MNRAS/358/1290 Mitronova S.: J/A+A/389/795 Mitronova S.N.: J/A+A/435/459 J/other/BSAO/57.5 J/PAZh/31/563 Mitsuda K.: J/ApJS/134/77 Mittaz J.P.D.: J/A+A/476/1191 J/MNRAS/274/1194 J/MNRAS/311/456 J/MNRAS/318/1073 Miura N.: J/other/PNAOJ/2.561 Miville-Deschenes M.-A.: II/243 Miyaji T.: J/ApJ/528/123 J/ApJS/161/21 J/PASJ/52/631 J/PASJ/52/997 J/PASJ/53/517 Miyama S.M.: J/PASJ/57/13 Miyamoto H.: J/other/APh/26.282 Miyamoto S.: IX/19 Miyauchi-Isobe N.: J/other/PNAOJ/5.75 J/other/PNAOJ/6.1 J/other/PNAOJ/6.107 J/other/PNAOJ/7.37 MIYAUCHI-ISOBE N.: VII/104 VII/135 VII/150 Miyauchi-Isobe N.: VII/179A Miyazaki A.: J/PASJ/56/261 J/PASJ/56/765 J/PASJ/58/529 Miyazaki M.: J/PASJ/55/1079

Miyazaki S.: J/AJ/127/180 J/ApJ/642/126 J/ApJS/118/177 J/ApJS/154/155 J/MNRAS/357/1357

J/PASJ/55/1079 J/PASJ/56/1025 J/PASJ/57/165 Miyazawa M.: J/ApJS/112/245 Miyoshi M.: J/PASJ/52/631 Mizumoto Y.: J/PASJ/57/165 Mizuno A.: J/ApJS/110/21 J/PASJ/51/791 J/PASJ/53/1081 J/PASJ/53/971 J/PASJ/57/917 Mizuno D.R.: J/AJ/124/2990 J/AJ/126/1423 V/114 Mizuno N.: J/PASJ/51/791 J/PASJ/53/971 Mizutani M.: J/A+A/423/579 M.J.: J/A+A/406/593 J/A+A/438/769 VIII/73 Mkrtichian D.E.: J/A+A/405/231 J/A+A/413/293 J/A+A/457/1033 J/MNRAS/353/310 Mobasher B.: J/A+A/363/517 J/AJ/131/1216 J/ApJ/607/665 J/ApJ/613/200 J/ApJ/624/135 J/ApJ/635/243

J/ApJS/137/279 J/ApJS/155/1 J/MNRAS/306/708 J/MNRAS/351/1290 J/MNRAS/352/1245 J/MNRAS/358/149 Mobasher: J/MNRAS/296/839 Mochejska B.J.: J/AcA/48/455 J/AJ/118/2211 J/AJ/118/346 J/AJ/121/2032 J/AJ/122/1383 J/AJ/122/2477

J/AJ/123/3460 J/AJ/124/1486 J/AJ/126/175 J/AJ/129/2856 MOCHNACKI S.: J/AJ/105/2291 Mochnacki S.: J/AJ/131/407 Mochnacki S.W.: J/AJ/130/767 Moehler S.: J/A+A/378/556 J/A+A/464/953 J/A+A/464/967 Moellenbrock G.A.: J/ApJ/616/110 J/ApJS/155/33 J/PASJ/52/997 Moellenhoff C.: J/A+A/368/16 J/A+A/398/49 J/A+A/422/907 Moeller C.S.: J/A+A/392/1 J/A+A/413/37 Moffat A.F.J.: J/A+A/331/949 J/A+A/463/981 J/AJ/109/817 J/AJ/120/2101 J/ApJS/93/187 J/MNRAS/324/33 MOFFAT A.F.J.: VII/5A Moffat J.W.: J/ApJ/636/721 Moffet A.T.: J/AJ/98/1148 Moffett A.J.: J/AJ/131/612 Moffett T. J.: J/PASP/111/812 Moffett T.J.: J/ApJS/117/135 Mohan R.: J/other/JApA/25.143 Mohan V.: J/A+A/265/32 J/A+A/372/71 J/A+AS/122/111 J/MNRAS/358/1290 J/other/BASI/29.519 Mohr J.J.: J/AJ/116/2108 J/AJ/120/2338 J/AJ/124/1266 J/ApJ/470/724 J/ApJ/639/816 Moiseev A.V.: J/A+A/470/605 J/AZh/82/420 Moitinho A.: B/ocl J/A+A/417/961 J/A+A/442/917 J/A+A/445/493 J/A+A/453/101 J/A+A/460/133

J/AJ/116/1801 J/MNRAS/367/1441 Moitinho: J/A+A/445/939 Mokhtarani A.: J/other/APh/26.282 Mokiem M.R.: J/A+A/437/467 MOLARO P.: J/A+A/290/885 Molaro P.: J/A+A/318/347 J/A+A/319/593 J/A+A/330/453 J/A+A/352/117 J/A+A/370/426 J/A+A/403/1105

J/A+A/416/1117 Molau S.: J/A+AS/125/399 Molendi S.: J/A+A/409/79 J/A+A/466/31 J/ApJ/514/148 J/ApJ/570/100 J/ApJS/99/701 J/MNRAS/327/771 Moles M.: J/A+A/285/404 J/A+A/310/31 J/A+A/310/8 J/A+A/387/26 J/A+A/393/389 J/A+A/416/475

J/A+A/420/873 J/A+A/470/39 J/A+AS/71/297 J/A+AS/91/285 J/ApJ/542/673 Molhant N.: II/230 Molina M.: J/ApJ/636/765 J/ApJS/170/175 Molinari E.: J/A+A/469/503 J/A+A/474/793 MOLINARI E.: J/A+AS/103/245 Molinari E.: J/A+AS/119/391 J/A+AS/140/155 J/MNRAS/246/576 Molinari S.: J/A+A/308/573 J/A+A/336/339 J/A+A/432/921 J/A+A/447/221 J/A+AS/101/59 Molkov S.: J/A+A/474/793 Molkov S.V.: J/PAZh/30/589 Molla M.: J/A+A/359/18 J/A+A/432/861 J/A+AS/130/513 J/MNRAS/358/521 Mollenhoff C.: J/A+A/415/63 Moller O.: J/MNRAS/345/1 Moller P.: J/A+A/457/79 Molnar L.: J/ApJ/481/95 MOLNAR L.A.: J/AJ/107/1433 Molster F.J.: J/A+A/380/609 J/A+A/382/184 J/A+AS/119/1 Molthagen K.: J/A+AS/126/509 J/A+AS/127/145 Momany I.: J/A+A/435/657

Momany Y.: J/A+A/379/436 J/A+A/390/931 J/A+A/413/343 J/A+A/440/901 J/A+A/444/831

J/A+A/455/271 J/A+A/464/927 J/A+A/464/953 J/A+A/464/967 J/A+A/467/1025 Momjian E.: J/AJ/130/2613 Monaco L.: J/A+A/464/201 J/A+A/465/815 J/AJ/124/3222 J/MNRAS/349/1278 Monbleau D.: J/AJ/131/2383 Monella R.: VII/103

Monet D.G.: I/263 I/268 I/283A I/284 I/293 I/297 J/AJ/103/638 J/AJ/124/1570 J/AJ/127/3034

J/ApJ/564/421 Moneti A.: II/243 J/A+A/331/70 J/A+AS/136/471 Monfils A. Houziaux L.: II/59B Monfils A.: II/86 MONFILS A.: III/39A V/38 Moni Bidin C.: J/A+A/466/181 J/A+A/466/931 MONIER R.: J/A+A/287/591 Moniez M.: J/A+A/303/137 J/A+A/389/149 J/A+A/421/509 J/A+A/454/185 MONIEZ M.: J/A+AS/109/447 Monin J.-L.: J/A+A/330/990 Monin J.-.L: J/A+A/425/927 Monin J-L: J/A+A/446/485 Monin J.-L.: J/A+A/468/353 Monk D.J.: J/MNRAS/234/583 Monnelly G.: J/other/GCN/1070.1 Monnet: J/A+AS/31/439 Monnier J.D.: J/ApJ/620/961 Monnier Ragaigne D.: J/A+A/408/465 Monnier-Ragaigne D.: J/A+A/365/1 J/A+A/405/99 J/A+A/408/67 Monnier-Ragnine D.: J/A+A/428/823 Monsignori Fossi B.C.: J/A+AS/125/149 Montagni F.: J/A+A/327/61 J/A+A/395/11 J/A+A/402/151 J/A+A/423/935 J/A+A/451/435 J/A+A/458/225

J/A+AS/130/109 Montalto M.: J/A+A/471/515 Montanari E.: J/ApJS/156/217 Montaruli T.: J/other/APh/26.282 Montegriffo P.: J/A+A/298/461 J/A+A/331/70 J/A+A/430/603 Monteiro M.J.P.F.G.: J/A+A/399/271 J/A+A/458/873 J/A+A/469/783 Montemayor T.J.: J/ApJS/156/227 Montenegro-Montes F.M.: J/AJ/129/1198 J/MNRAS/370/1034 Montes D.: J/A+A/312/221 J/A+A/379/976 J/A+AS/123/473 J/A+AS/128/485 J/A+AS/146/103

J/ApJS/123/283 J/MNRAS/328/45 Montesinos B.: J/A+A/365/110 J/A+A/378/116 J/A+A/379/564 Montgomery K.A.: J/AJ/106/181 J/AJ/108/585 Montignac G.: I/272 J/A+AS/125/399

Montmerle T.: J/A+A/350/883 J/A+A/357/219 J/A+A/359/113 J/A+A/372/173 J/A+A/420/945

J/A+A/421/623 J/A+A/429/963 J/A+A/468/353 Montoya C.: J/ApJ/621/104 Moody J.W.: J/AJ/109/2376 J/AJ/121/66 J/AJ/123/1292 J/AJ/127/1943 J/AJ/130/2571 J/AJ/95/1629 Moody K.: J/ApJ/614/167 Moody R.A.: J/A+A/366/481 Moody S.: J/MNRAS/329/87 J/MNRAS/352/939 Mookerjea B.: J/A+A/426/119 J/MNRAS/367/1609 Mooney C.J.: J/A+A/419/1123 Moor A.: J/A+A/417/745 J/A+A/418/131 J/A+A/463/1227 Moore A.W.: J/ApJS/155/257 MOORE Ch.E.: VI/26 Moore D.C.: J/ApJS/162/1 Moore S.A.W.: J/AJ/128/1558 Moore T.J.T.: J/A+A/461/11 J/A+A/474/891 J/MNRAS/366/1096 J/MNRAS/379/663 Moorwood A.: J/A+A/382/60 J/AJ/120/2206 J/AJ/122/2205 J/AJ/125/1107 J/ApJ/604/521 Moorwood A.F.M.: J/A+A/372/364 J/A+A/398/49 J/A+A/418/465 MOORWOOD A.F.M.: J/A+AS/108/621 Moorwood A.F.M.: J/AJ/131/1891 Moos H.W.: J/ApJS/146/1 J/ApJS/146/125 Mora A.: J/A+A/365/110 J/A+A/378/116 J/A+A/379/564 Mora M.D.: J/A+A/464/495 MORABITO D.: I/77 Morabito D.D.: I/76 Morabito L.: I/152 Morale F.: J/A+A/341/L79 Moralejo A.: J/A+A/390/39 J/A+A/406/L9 J/A+A/410/813 Morales J.C.: J/A+A/478/507 Morales-Rueda L.: J/A+A/466/1031 Moran E.C.: J/AJ/124/3465 J/AJ/131/1948 J/ApJ/461/127 J/ApJ/632/799 Moran J.M.: J/ApJ/393/149 J/ApJ/615/702 Moran L.: J/ApJ/636/765 Moran S.M.: J/ApJ/634/977 Morand F.: J/A+A/312/879 Morando B.: J/A+AS/125/399 Moraru D.: J/AJ/131/527 Moraux E.: J/A+A/400/891 J/A+A/448/189 J/A+A/471/499 J/MNRAS/370/954 J/MNRAS/377/741 Morbidelli L.: J/A+A/410/905 J/MNRAS/349/793 Morbidelli R.: J/A+A/448/579 Moreau C.: J/A+A/428/1043 J/A+A/451/1159 Moreau O.: J/A+A/303/137 MOREAU O.: J/A+AS/109/447 Moreau O.: J/A+AS/111/169 J/A+AS/123/529 Morel M.: II/7A Morel T.: J/A+A/412/495 J/A+A/423/677 J/A+A/426/1007 J/A+A/457/651 J/MNRAS/351/1290 Morell N.I.: B/sb9 V/122 Morelli L.: J/A+A/416/507 Morelli P.L.: J/A+A/314/541 Moreno R.: J/A+AS/122/271 Moreno-Corral M.A.: J/A+AS/145/187 J/other/RMxAA/38.141 Moretti A.: IX/34 J/A+A/435/799 J/A+A/448/L9 J/A+A/452/179 J/A+A/474/793 J/A+AS/140/155 Morey A.: J/other/APh/26.282 Morgan D.: J/A+AS/113/539 Morgan D.H.: J/A+A/369/932 J/A+AS/122/507 J/ApJS/83/87 J/MNRAS/341/534 J/MNRAS/355/1196

J/MNRAS/362/689 V/127 VII/183 Morgan E.: J/other/GCN/1070.1 Morgan H.R.: I/40 MORGAN H.R.: I/80 Morgan J. J.: J/ApJS/121/171 Morganti R. Garrett M.A.: J/A+A/424/371

Morganti R.: J/A+A/385/39 J/A+A/439/487 J/A+A/453/493 J/AJ/123/1784 J/MNRAS/285/898

J/MNRAS/324/1131 MORGULEFF N.: II/26A Moriarty-Schieven G.: J/ApJ/559/307 Morin D.: I/196 MORIN D.: IV/12 IV/4 V/3 Moriondo G.: J/A+AS/137/101 J/ApJ/510/637 Morisaki S.: J/PASJ/52/631 MORISSET C.: J/A+A/303/175 Morisset C.: J/AJ/130/1707 Morita M.: J/ApJ/571/665 J/ApJS/140/143 Morlet G.: J/A+A/369/552 J/A+A/396/933 J/A+AS/135/499 J/A+AS/145/67 Morley J.E.: J/MNRAS/326/1161 Moro D.: J/A+A/370/503 Morokuma T.: J/PASJ/57/165 Moro-Martin A.: J/PASP/118/1690 Morossi C.: J/A+AS/126/267 J/ApJ/601/485 J/ApJ/613/312 J/ApJ/634/1319 J/PASP/113/1365 Morras: J/A+AS/142/25 Morras R.: VIII/75 VIII/76 Morrell N.: III/204 J/A+A/378/861 J/A+AS/118/231 J/ApJ/490/328 Morrell N.I.: J/ApJS/99/135 J/MNRAS/366/739 Morrill A.C.: J/ApJS/115/185 Morris D.: J/A+A/448/L9 J/A+AS/143/145 Morris D.C.: J/A+A/474/793 Morris D.J.: J/MNRAS/335/275 J/MNRAS/342/1299 Morris M.: J/A+A/314/871 J/A+A/426/81 J/AJ/130/1763 J/ApJ/581/258 Morris M.R.: J/ApJ/589/225 J/ApJ/613/1179 Morris P.: J/A+A/384/473 J/PASP/118/1690 Morris R.A.H: J/A+A/476/59 Morris R.A.H.: J/MNRAS/362/689 V/127 Morris S.: J/ApJ/471/694 Morris S.L.: J/ApJ/565/208 J/ApJS/113/1 J/ApJS/116/211 J/ApJS/116/247 J/ApJS/129/475 J/MNRAS/318/703

J/MNRAS/367/1261 Morrison G.: J/A+A/451/57 Morrison G.E.: J/AJ/126/119 J/AJ/129/26 J/AJ/129/31 J/ApJ/581/844 J/ApJS/146/267 Morrison H.: J/AJ/120/2496 Morrison H.L.: J/AJ/100/1191 J/AJ/125/2502 J/ApJ/616/804 Morrison J.: J/ApJS/154/155 Morrison J.E.: J/AJ/131/250 Morrison L.V.: I/99 J/A+A/312/1031 Morrison N.D.: J/A+A/408/305 Morrissey P.: J/AJ/130/825 J/ApJ/619/L95 Morrissey P.F.: J/ApJS/164/38 Morse J.A.: J/AJ/124/1144 J/ApJ/613/343 Morse R.: J/other/APh/26.282 Morshidi-Esslinger Z.: J/MNRAS/304/297 Mortier A.M.J.: J/MNRAS/372/1621 Morton D.C.: J/ApJS/147/61 MORTON M.D.: J/AJ/105/2265 Mortsell E.: J/ApJ/561/106 Moscadelli L.: J/A+A/417/615 Moscardini L.: J/MNRAS/312/540 Moscoso L.: J/A+A/303/137 MOSCOSO L.: J/A+AS/109/447 Moshir M.: II/156A J/AJ/131/250 Moshir M.M.: V/107 Moshkalev V.G.: J/AZh/72/80 Moskalik P.: J/A+A/341/818 Mosoni L.: J/ApJS/155/33 Moss C.: J/MNRAS/357/1337 Moss D.: J/MNRAS/362/1371 J/MNRAS/369/156 J/MNRAS/378/995 Mossakovskaya L.V.: III/208 Mosser B.: J/A+A/478/197 Mossman A.: J/ApJ/618/123 J/ApJ/644/829 J/ApJ/645/955 Mossman A.E.: J/AJ/131/527 Motch C.: J/A+A/235/185 J/A+A/318/111 J/A+A/476/1191 J/A+AS/122/201 J/A+AS/132/341 Mothe-Diniz T.: J/A+AS/134/463 Motohara K.: J/AJ/131/1608 J/ApJ/648/7 J/PASJ/53/653 J/PASJ/57/165 J/PASJ/58/313

Motte F.: J/A+A/372/L41 J/A+A/421/623 J/A+A/458/789 J/A+A/472/519 J/A+A/476/1243

J/ApJS/156/127 Mottola S.: J/A+AS/113/115 Mottram J.C.: J/A+A/461/11 MOUGINS B.: I/74 Mouhcine M.: J/A+A/393/167 J/A+A/402/425 J/MNRAS/362/1143 VI/125 Mouillet D.: J/A+A/475/1053 Moujtahid A.: J/A+AS/129/289 Mould J.: J/ApJ/348/98 J/ApJS/142/71 J/ApJS/154/623 Mould J .R.: J/ApJ/523/540 Mould J. R.: J/PASP/108/682

Mould J.R.: J/AJ/113/1001 J/AJ/117/1700 J/AJ/123/873 J/AJ/124/1954 J/AJ/124/690 J/AJ/128/16 J/ApJ/297/582 J/ApJ/298/544 J/ApJ/463/26 J/ApJ/507/655 J/ApJ/508/491 J/ApJ/512/48 J/ApJ/514/614 J/ApJ/521/155 J/ApJ/529/723

Mould J.R. McCarthy J.K.: J/AJ/110/2319 MoulD J.R.: VII/75 Moultaka J.: J/A+A/433/117 Mourard D.: J/A+A/312/879 J/A+A/453/309 Mourard: J/A+A/428/587 Moustakas J.: J/ApJS/164/81

Moustakas L.A.: J/A+A/434/53 J/A+A/454/423 J/A+A/478/83 J/ApJ/475/445 J/ApJ/600/L155

J/ApJ/600/L99 J/ApJ/613/200 Moutou C.: J/A+A/424/L31 J/A+A/466/743 J/A+A/476/1347 Movit S.M.: J/other/APh/26.282 Movsessian T.A.: J/A+A/470/605 Mowlavi N.: J/A+A/360/509 J/A+A/374/968 J/A+A/411/L59 J/A+AS/128/471 J/A+AS/135/405 V/117 Moy E.: J/A+A/403/493 Moya A.: J/A+A/426/247 J/A+A/450/715 J/A+A/456/261 J/A+A/472/565 Mozurkewich D.: J/AJ/114/2822 Msezane A.Z.: J/A+A/410/349 J/A+A/473/995 J/ApJS/121/265 Mucciarelli A.: J/A+A/462/599 Much R.: J/A+AS/143/145 J/ApJ/607/L33 Mucke A.: J/ApJS/123/79 Mueller H.: J/A+A/391/547 Mueller H.S.P.: J/A+A/428/1019 Mueller K.R.: J/A+AS/140/327 Mueller T.G.: J/A+A/338/340 Mueller V.: J/A+A/336/35 J/A+A/459/L1 Muena C.: J/AJ/112/2408 Muench A.A.: J/AJ/128/1254 J/AJ/131/1574 J/AJ/131/527 J/ApJ/573/366 J/ApJ/593/1093 J/ApJS/160/319 MUEYESSEROGLU Z.: J/A+AS/106/373 Mufson S.: J/AJ/127/1158 Mufson S.L.: II/266 Mujica R.: J/A+A/349/619 J/A+A/414/487 J/A+A/433/151 J/ApJS/117/319 J/MNRAS/351/1290 Mujica R.M.: J/AJ/124/1283 Mujunen A.: J/A+AS/132/305 Mukadam A.S.: J/ApJ/607/426 Mukai K.: J/ApJ/561/189 Mukai T.: J/PASJ/58/L29 Mukanov D.B.: J/A+AS/147/5 Mukherjee R.: J/ApJS/101/259 J/ApJS/107/227 J/ApJS/123/79 Mukherjee S.: J/ApJ/508/314 Mulas G.: J/A+A/413/343 Mulchaey J.: J/ApJ/617/718 Mulchaey J.S.: J/ApJ/496/39 J/ApJ/539/136 J/ApJ/549/172 J/ApJ/643/680 J/ApJS/146/353 Muller E.: J/AJ/130/2613 J/MNRAS/339/105 MULLER P.: J/A+AS/102/643 Muller P.: J/A+AS/126/273 Muller V.: J/A+A/462/811 J/AJ/122/2222 J/AJ/123/51 Mullis C.R.: J/A+A/461/977 J/ApJ/580/774 J/ApJ/594/154 J/ApJ/617/192 J/ApJS/149/29 J/ApJS/162/304 Muminov M.: V/27 Munari U.: III/238 J/A+A/311/484 J/A+A/323/113 J/A+A/343/806 J/A+A/370/503 Munari U: J/A+A/372/145

Munari U.: J/A+A/383/188 J/A+A/386/237 J/A+A/401/781 J/A+A/406/995 J/A+A/415/145 J/A+A/417/1055 J/A+A/418/869 J/A+A/434/1107 J/A+A/452/567 J/A+A/465/481 J/A+AS/116/1 J/A+AS/137/521 J/A+AS/143/343 J/A+AS/146/407

MUNARI U.: J/A+AS/93/383 Munari U.: J/MNRAS/327/23 Munch W.: J/A+AS/75/67

Mundt R.: J/A+A/312/439 J/A+A/396/513 J/A+A/416/555 J/A+A/417/557 J/A+A/430/1005

J/A+AS/114/109 Muneer S.: J/A+A/469/713 J/MNRAS/372/1315 Mungall F.: J/AJ/126/3007 J/AJ/128/463 Munich K.: J/other/APh/26.282

Munn J.A.: J/AJ/121/2308 J/AJ/121/2331 J/AJ/121/31 J/AJ/122/503 J/AJ/123/567 J/AJ/125/1817 J/AJ/126/1023 J/AJ/126/2209 J/AJ/126/2579 J/AJ/127/3034 J/AJ/127/3553 J/AJ/128/426 J/AJ/129/2542 J/AJ/131/582 J/ApJ/607/426 J/ApJS/109/45 J/ApJS/155/257

Muno M.P.: J/ApJ/589/225 J/ApJ/613/1179 J/ApJS/165/173 Munoz J.A.: J/AJ/131/527 Munoz R.R.: J/AJ/130/2677 Munoz-Arjonilla A.J.: J/A+A/452/739 Munoz-Tunon C.: J/A+A/471/17 J/A+AS/129/455 J/AJ/110/212 J/ApJ/619/270 Murakami H.: J/A+A/376/997 J/A+A/403/943 J/PASP/117/199 Murakami I.: J/A+A/460/331 Murakami T.: J/ApJ/609/935 Murakawa K.: J/A+A/470/957 Muraki Y.: J/MNRAS/330/137 J/MNRAS/331/609 Murata Y.: J/ApJ/616/110 J/ApJS/155/33 J/PASJ/52/997 Muratorio G.: J/A+A/258/423 J/A+AS/136/95 Murayama T.: J/ApJ/648/7 J/PASJ/57/165 J/PASJ/57/881 J/PASJ/58/313 Murdock T.L.: II/94 V/107 Muriel H.: J/A+A/445/819 J/AJ/110/1032 J/AJ/129/1237 J/AJ/130/2012 Murowinski R.: J/AJ/127/2455 Murozono K.: J/ApJ/648/7 Murphy B. W.: J/AJ/115/708 Murphy D.C.: J/AJ/128/2239 VIII/39 Murphy D.W.: J/PASJ/52/997 Murphy E.J.: J/ApJ/633/871 Murphy E.M.: J/ApJS/105/369 J/ApJS/140/331 J/ApJS/146/1 Murphy H.P.: J/MNRAS/211/637 Murphy J.: J/MNRAS/379/289 MURPHY J.K.: I/73 Murphy J.K.: I/90 Murphy M.T.: J/MNRAS/311/741 J/MNRAS/345/609 J/MNRAS/354/L25 J/MNRAS/354/L31 J/MNRAS/371/495

J/other/PASA/19.455 Murphy R. J.: J/ApJS/120/409 Murphy T.: J/A+A/392/795 J/MNRAS/329/227 J/MNRAS/371/898 J/MNRAS/377/787 VIII/70A VIII/78

VIII/81 VIII/82 Murphy T.W.JR: J/AJ/121/97 MURRAY C.A.: I/129 Murray C.A.: I/85 MURRAY C.A.: J/AJ/103/514 Murray C.D.: J/A+AS/121/65 J/A+AS/136/257 Murray M.A.: J/AJ/96/1655 Murray S.: IX/3 J/AJ/115/592 MURRAY S.: VII/18

Murray S.S.: IX/22 J/A+A/460/133 J/ApJ/560/675 J/ApJ/577/738 J/ApJ/582/382 J/ApJ/588/1009 J/ApJ/599/1067 J/ApJ/606/466 J/ApJ/608/781 J/ApJ/609/735 J/ApJ/610/247 J/ApJ/635/1053 J/ApJ/641/140 J/ApJS/160/319 J/ApJS/161/9 J/other/Sci/292.2290

Murtagh F.: J/ApJ/508/314 VI/37 Murthy J.: II/269 Musaev F.: J/A+A/430/255 J/MNRAS/359/865 Musaev F.A.: J/A+A/414/949 J/A+AS/136/453 J/ApJ/634/451 J/AZh/81/541

Musella I.: J/A+A/315/348 J/A+A/383/398 J/A+A/428/339 J/A+A/462/599 J/MNRAS/365/1357

J/MNRAS/371/1503 Musgrove A.: VI/73 VI/74

Mushotzky R.F.: J/AJ/128/1483 J/AJ/128/1501 J/ApJ/253/485 J/ApJ/633/L77 J/ApJS/96/303

J/MNRAS/356/568 MUSIELAK G.: III/120 Musser J.: J/AJ/127/1158 Mutafov A.S.: VI/90 Mutchler M.: J/AJ/120/2747 J/ApJ/633/871 Muterspaugh M.W.: J/AJ/130/2866 J/ApJ/636/1020 Muthu C.: J/AJ/126/2963 Mutoh H.: J/ApJ/382/636 Mutschke H.: J/A+A/423/983 J/A+A/449/583 Muxlow T.W.B: B/avo.rad Muxlow T.W.B.: J/A+A/472/805 J/AJ/130/1977 J/MNRAS/358/1159 J/MNRAS/376/371 Muzerolle J.: J/AJ/128/805 J/AJ/131/1574 J/AJ/131/527 J/ApJ/620/1010 J/ApJ/638/897 Muzzin A.: J/AJ/122/2587 J/AJ/127/334 Muzzio J.C.: J/AJ/82/345 J/AJ/82/474 MYAKUTIN V.I.: V/43 Myers J.R.: V/102 V/105 V/109 V/95 Myers P.: J/AJ/131/1574 Myers P.C.: J/A+A/413/1177 J/AJ/126/1916 J/ApJ/526/788 J/ApJ/566/993 J/ApJS/123/233 J/ApJS/125/161

J/ApJS/95/457 Myers S.T.: J/AJ/130/923 VIII/72 Myznikov A.: I/302

N.: J/A+A/428/587 J/MNRAS/326/959

Nadeau D.: J/A+A/323/139 J/A+A/330/990 J/A+A/344/943 J/A+A/360/1107 J/A+A/394/225

J/MNRAS/317/234 Nadzhip A.E.: J/AZh/71/618 Naef D.: J/A+A/356/590 J/A+A/375/205 J/A+A/375/L27 J/A+A/379/999 J/A+A/388/632 J/A+A/390/267

J/A+A/392/215 J/A+A/406/373 J/A+A/410/1039 J/A+A/410/1051 J/A+A/414/351 J/A+A/415/391 Nagao T.: J/A+A/459/85 J/ApJ/648/7 J/PASJ/57/165 J/PASJ/57/881 J/PASJ/58/313 Nagar N.M.: J/A+A/409/115 J/A+A/435/521 J/AJ/120/1364 Nagashima C.: J/ApJ/596/1064 J/MNRAS/353/705 J/MNRAS/354/980 J/MNRAS/368/534 J/PASJ/55/1079

Nagata T.: J/AJ/117/446 J/ApJ/596/1064 J/MNRAS/353/705 J/MNRAS/354/980 J/MNRAS/368/534

J/PASJ/55/1079 Nagayama T.: J/ApJ/596/1064 J/MNRAS/353/705 J/MNRAS/354/980 J/MNRAS/368/534 J/PASJ/55/1079 Nagy A.: J/A+A/454/257 NAGY T.A. et al.: II/108 NAGY T.A.: I/133 Nagy T.A.: II/71 II/96 IV/13 Nahar S.: J/A+A/389/716 Nahar S N: J/A+AS/135/347

Nahar S.N.: J/A+A/293/967 J/A+A/408/789 J/A+A/413/779 J/A+A/437/345 J/A+A/448/779

J/A+A/457/721 J/A+AS/119/509 J/A+AS/144/141 J/A+AS/147/253 J/ApJS/167/315 Nahnhauer R.: J/other/APh/26.282 Naidu B.N.: J/other/BASI/29.519 Naim A.: J/MNRAS/274/1107 Nair P.: J/ApJ/588/218 Najarro F.: J/A+A/437/467 J/ApJ/581/258 Najita J.: J/ApJS/161/9 J/PASP/118/1690 Najita J.R.: J/ApJ/541/977 J/ApJ/641/140 NAKADA N.: VII/140

Nakada Y.: J/A+A/385/884 J/ApJS/106/463 J/ApJS/155/401 J/MNRAS/353/705 J/MNRAS/364/117

J/MNRAS/370/1979 J/PASJ/56/261 J/PASJ/56/765 J/PASJ/58/529 Nakagawa M.: J/ApJS/63/661 J/ApJS/74/291 J/PASJ/57/917 Nakai T.: J/PASJ/56/295 Nakajima H.: J/PASJ/55/635 J/PASJ/55/653 Nakajima J.: J/PASJ/52/631 J/PASJ/52/997 Nakajima T.: J/AJ/121/2189 Nakajima Y.: J/AJ/131/1608 J/ApJ/596/1064 J/MNRAS/353/705 J/MNRAS/354/980 J/PASJ/55/1079 Nakamura F.E.: J/AJ/111/42 Nakamura M.: J/PASJ/52/631 Nakamura O.: J/MNRAS/366/144 Nakamura T.: J/ApJ/609/935 J/MNRAS/330/137 Nakanishi K.: J/AJ/111/42 J/ApJS/112/245 NAKANO M.: III/177 III/177 III/177 III/177 Nakashima J.: J/A+A/470/957 J/PASJ/53/293 J/PASJ/54/719 J/PASJ/55/203 J/PASJ/55/229 J/PASJ/56/261 Nakashima J.-I.: J/PASJ/56/765 Nakata F.: J/MNRAS/357/1357 J/PASJ/55/1079 Nakaya H.: J/MNRAS/353/705 J/MNRAS/354/980 J/MNRAS/368/534 J/MNRAS/370/1979 J/PASJ/55/1079 Nakazaki S.: J/A+A/436/1141 Nakos T.: J/A+A/402/1157 J/A+AS/124/353 Nalezyty M.: J/A+A/420/507 J/A+AS/125/337 Nam J.W.: J/other/APh/26.282 Nan R.: VIII/44 Nan R.D.: J/other/AcApS/16.103 Nanda Kumar M.S.: J/MNRAS/353/1025

Nandra K.: J/AJ/129/2074 J/ApJ/582/615 J/ApJ/642/126 J/MNRAS/350/785 J/MNRAS/356/568

J/MNRAS/371/221 NANDY K.: II/130 Nandy K. Jamar C.: II/59B NANDY K.: V/38 Nankivell G.R.: J/MNRAS/330/137 Nantais J.B.: J/AJ/131/1416 Naoi T.: J/A+A/449/583 J/MNRAS/370/1979 Napier M.G.: J/AJ/121/2148 Napier W.M.: J/A+A/310/353 Napiwotzki R.: J/A+A/378/556 J/A+A/400/877 J/A+A/400/939 J/A+A/430/223 J/A+A/447/173 Napolitano N.: J/A+A/428/339 Napolitano N.R.: J/MNRAS/369/120 Naranjo O.: J/AJ/127/1158 J/AJ/131/527 Narashimha D.: J/A+A/402/113 Narayanan V.: J/AJ/123/3409 Narbutis D.: J/PASJ/56/1025 Nardetto: J/A+A/428/587 Nardetto N.: J/A+A/453/309 J/A+A/476/73 Nardo G.: VII/69 Nariai K.: J/ApJ/648/7 J/PASJ/57/165 NASA / ESA / SERC: VI/27 VI/75 NASA / ESA: VI/110 VI/99 Naseri S.: J/A+A/475/1053 Nasi E: J/A+AS/106/275 Nasonov D.S.: III/243 III/245 Nass P.: J/A+AS/127/145 Nassau J.J.: II/221A III/15 III/15 III/15 III/76A III/76A III/76A J/ApJ/120/129 J/ApJ/125/195 Nassiopoulos G.: J/ApJS/99/223 Natale E.: J/A+A/368/845 NATALE V.: J/A+A/284/233 J/A+A/291/261 Natale V.: J/A+AS/101/153 Natarajan P.: J/ApJ/591/53 J/ApJ/640/603 National Geographic Society: VI/25 NATIONS H.L.: J/A+AS/103/183 Nations H.L.: J/A+AS/139/513 Natta A.: J/A+A/452/245 NAUBER U.: VII/78 VIII/5 Naugolnaya M.N.: J/AZh/71/684 J/AZh/72/613 Naugol’naya M.N.: J/AZh/73/835 Naugolnaya M.N.: J/other/BSAO/40.5 NAUMANN M.: II/200 Naumann M.: J/A+AS/93/463 NAUMOVA A.A.: I/43 Naumova A.A.: I/44 I/82 Naundorf C.E.: J/ApJ/439/983 Navarrete M.: J/AJ/112/2408 Navasardyan H.: J/A+A/452/567 J/MNRAS/328/1181 J/MNRAS/369/1880 NAVE G.: J/A+AS/102/269 Nave G.: J/ApJS/94/221 Nayakshin S.: J/ApJ/643/1011 Naylor T.: J/A+A/384/504 J/MNRAS/274/1194 J/MNRAS/335/291 J/MNRAS/341/805 J/MNRAS/346/1143

J/MNRAS/351/1401 J/MNRAS/356/1583 J/MNRAS/356/89 J/MNRAS/358/341 J/MNRAS/361/1091 J/MNRAS/375/1220 J/MNRAS/376/580

Naze N.: J/MNRAS/372/661 Naze Y.: J/A+A/395/499 J/A+A/418/841 J/A+A/463/981 J/MNRAS/375/145 N.C.: J/A+A/404/775 Neale L.T.: J/A+AS/133/257 Neckel Th.: II/34A II/62 Nedialkov P.: J/A+A/389/439 Nedialkov P.L.: J/A+A/466/501 Neely R.K.: J/AJ/119/1793 Neeser M.J.: J/MNRAS/329/877 NEFEDIAVA A.I.: I/17 Nefedjev J.: J/A+A/444/625 Neff J.E.: J/A+A/310/173 Neff S.G.: J/AJ/113/1011 J/AJ/116/180 J/AJ/130/825 J/ApJ/619/L95 J/ApJS/164/38 Neger T.: J/A+A/388/704 Negoro H.: IX/19 Negueruela I.: J/A+A/461/631 J/A+A/462/1081 J/A+A/476/1191 Neill J.D.: J/AJ/111/1596 J/AJ/114/1471 J/ApJ/339/53 Neill Reid I.: J/MNRAS/318/1206 Neilsen E.H.: J/A+A/444/643 J/AJ/123/3409 J/AJ/124/1651 J/ApJ/483/745 J/PASJ/55/771 Neilsen E.H.JR: J/AJ/127/2210 Neilsen E.H.Jr: J/AJ/131/1674 Neilsen E.H.JR: J/AJ/132/676

Neiner C.: J/A+A/403/659 J/A+A/409/275 J/A+A/411/565 J/A+A/451/1053 J/A+A/457/651

J/A+A/462/683 J/A+A/472/565 J/A+A/472/577 NEIZVESTNAYA O.M.: VII/173 Neizvestny S.: IX/32 NEIZVESTNY S.I.: VII/173 Nelan J.E.: J/AJ/128/1558 Nelemans G.: J/A+A/378/556 J/A+A/466/1031 Nelson A.: J/ApJS/137/117 Nelson A.E.: J/ApJ/563/629 J/ApJ/566/103 Nelson B.: J/ApJ/510/659 Nelson B. O.: J/ApJS/125/349 Nelson B.O.: J/AJ/126/63 J/AJ/127/646 J/ApJ/569/23 Nelson C.: J/ApJ/542/257 J/ApJ/552/289 J/ApJ/598/597 Nelson C.A.: J/AJ/127/334 J/ApJ/562/337 Nelson C.H.: J/AJ/125/1696 J/AJ/127/1360 NELSON E.: V/71 Nelson R.H.: J/AJ/110/2400 J/other/IBVS/5493 Nelson-Patel K.: J/AJ/131/527 Nemec J.: J/AJ/115/1840 J/AJ/115/1856 Nemec J. M.: J/AJ/117/286 NEMEC J.M.: J/AJ/103/460 Nemec J.M.: J/AJ/108/222 J/AJ/109/618 J/AJ/110/1186 J/ApJ/336/780 Nemeth P.: I/279 Neri R: B/iram Neri R.: J/A+A/475/559 Nesci R.: J/A+A/327/61 J/A+A/395/11 J/A+A/402/151 J/A+A/451/435 J/A+AS/130/109 Nesci R. Sclavi S.: J/A+A/423/935 Neslusan L.: J/other/CoSka/33.159 NESTEROV V.V.: I/159A Nesterov V.V.: III/182 Nestor D.B.: J/ApJ/636/610 Netherlands Foundation for Research in Astronomy ((NFRA/ASTRON): VIII/62 Netolicky M.: J/A+A/459/849

Netopil M.: J/A+A/443/157 J/A+A/454/171 J/A+A/454/179 J/A+A/459/871 J/A+A/462/157

J/A+A/462/591 Netzer H.: J/A+A/369/57 J/ApJ/613/682 Neuforge-Verheecke C.: J/A+A/328/261 Neugebauer G.: II/2B J/ApJS/127/1 NEUGEBAUER G.: VII/109

Neuhaeuser R.: J/A+A/297/391 J/A+A/319/184 J/A+A/320/185 J/A+A/341/163 J/A+A/353/1044 J/A+A/359/227 J/A+A/423/1029 J/A+A/439/1149 J/A+AS/123/329 J/A+AS/124/449 J/A+AS/146/323 J/PASJ/53/1081

Neuhauser R.: J/A+A/312/439 J/A+A/325/647 J/A+A/468/1039 Neumann D.: J/ApJ/514/148 Neumann D.M.: J/A+A/425/367 Neumann M.: J/A+AS/106/303 NEUPERT W.M.: J/ApJS/91/461 Neuschaefer L.: J/AJ/116/102 Newberg H.J.: J/AJ/121/2308 J/AJ/123/567 J/AJ/126/2209 J/AJ/126/2579 J/ApJ/588/824 J/ApJS/123/377

J/ApJS/145/245 Newberry M.V.: J/AJ/109/2376 Newcomb M.: J/AJ/121/2308 J/AJ/121/2331 J/AJ/123/567 Newcomer R.E.: II/269 Newell B.: VII/28 Newman J.: J/ApJ/633/174 Newman J.A.: J/AJ/121/31 J/AJ/127/3121 J/ApJ/625/6 J/MNRAS/371/221 Newman P.R.: J/AJ/123/3409 J/PASJ/55/771 Newmark J.: J/ApJ/500/554 Newsam A.M.: J/MNRAS/342/802 J/MNRAS/362/1371 Newton L.M.: J/other/PASAu/6.471 Newton-McGee K.J.: J/ApJ/634/L89 Ng Y.K.: J/A+A/329/943 J/A+A/338/581 J/A+A/392/1129 J/A+AS/123/115 Ngeow C.: J/A+A/411/361 J/MNRAS/369/723 Niarchos P.: J/A+A/373/950 J/A+A/374/1056 J/A+A/374/164 Niarchos P.G.: J/A+A/369/960 J/A+AS/124/291 Niarchos P.G: J/A+AS/133/7 Nicastro L.: J/A+A/455/813 J/A+A/469/503 J/A+A/474/793 Nice D.J.: J/ApJ/605/759 Nichol R.: J/AJ/124/1651 J/AJ/125/1817 J/MNRAS/360/587 J/MNRAS/372/425 J/MNRAS/375/68

Nichol R.C.: J/AJ/120/1 J/AJ/121/2308 J/AJ/121/2331 J/AJ/122/503 J/AJ/123/1807 J/AJ/123/567 J/AJ/126/1023 J/AJ/126/1720 J/AJ/126/2209 J/AJ/126/2579 J/AJ/130/968 J/AJ/131/1288 J/AJ/131/2018 J/ApJS/126/209 J/ApJS/148/243 J/ApJS/155/257 J/MNRAS/274/1071 J/MNRAS/341/1093 J/PASJ/55/771

NICHOL R.C.: VII/175 Nicholas J.P.: J/A+A/467/785 Nichols J.: J/A+A/411/565 J/ApJ/589/509 Nichols J.S.: III/188 J/A+AS/105/305 J/A+AS/116/257 Nichols-Bohlin J.: III/115 Nicholson K.L.: J/MNRAS/311/456 Nicholson W.: I/85 Nicklas H.: J/A+A/398/49 Nicolaci da Costa L.: J/AJ/105/1637 Nicolet B.: I/196 II/27 II/35 II/48 II/51 NICOLET B.: J/A+A/299/557 Nicolet B.: J/A+A/398/1073 J/A+A/406/373 NICOLET B.: J/A+AS/113/483 Nicolet B.: J/A+AS/122/51 J/A+AS/125/399 NICOLET B.: V/24 Nicolson G.: J/AJ/110/880 J/PASJ/52/997 Nicolson G.D.: J/AJ/127/3609 J/AJ/130/2529 Nicotra M.A.: J/AJ/110/1032 Nidever D.L.: J/ApJS/141/503 Niederste-Ostholt M.: J/AJ/126/2579 NIEDZIELSKI A.: J/A+AS/108/669 NIEHUES M.: II/200 Nielbock M.: J/A+A/426/119 NIELL A.E.: I/169 Nielsen K.E.: J/ApJS/157/138 Niemczura E.: J/A+A/404/689 J/A+A/415/241 Niemela T.: J/A+A/427/769 Niemela V.S.: J/AJ/120/2101 J/MNRAS/324/33 NIEMELA V.S.: J/PASP/98/1133 Nieppola E.: J/A+A/445/441 J/AJ/133/1947 Niessen P.: J/other/APh/26.282 NIETO J.-L.: J/A+AS/103/573 Nieto J.-L.: J/A+AS/116/429 NIEUWENHUIJZEN H.: III/141 Niezgoda J.: J/A+AS/135/273 Ni-Ino M.: J/other/PNAOJ/2.561 Nikanorova I.N.: III/201 J/PAZh/32/672 Niklas S.: J/A+AS/114/21 Nikolaev S.: J/A+A/411/361 Nikolashvili M.G.: J/A+A/376/775 J/A+A/402/151 J/other/Ap/31.714 NIKOLOFF I.: I/97 Nikula P.O.: J/A+A/350/513 Nilakshi: J/A+A/374/504 Nilakshi N.: J/A+A/381/65 Nilson P.: VII/26D Nilsson A.: J/A+A/358/257 J/A+AS/144/441 Nilsson K.: J/A+AS/116/157 J/A+AS/130/109 J/A+AS/132/305 J/A+AS/132/31 J/A+AS/143/357 NILSSON K.: J/ApJ/413/453 Nilsson K. Pursimo T.: J/A+AS/129/577 Ninkov Z.: J/A+AS/119/221 J/AJ/117/548 J/AJ/120/2638 Nisenson P.: J/A+A/410/1051 J/A+AS/134/545 NISHIDA S.: VII/140 Nishihara E.: J/ApJ/635/214 Nishikawa A.: J/ApJS/102/411 Nishimura M.: J/A+A/420/673 Nishio M.: J/PASJ/52/631 Nishiuchi M.: J/PASJ/55/161 Nishiyama S.: J/MNRAS/368/534 J/MNRAS/370/1979 Nisini B.: J/A+A/419/999 J/A+A/422/141 Nissen H.D.: J/A+A/466/949 NISSEN P.E.: II/180

Nissen P.E.: J/A+A/275/101 J/A+A/326/751 J/A+A/371/943 J/A+A/390/235 J/A+A/415/993 J/A+A/422/527 J/A+A/440/321 J/A+AS/102/603 J/A+AS/117/317 J/A+AS/141/491 Nitschelm B.: J/A+AS/125/399 NITSCHELM C.: II/163 Nitschelm C.: II/164 NITSCHELM C.: II/165 II/166 II/167 Nitschelm C.: J/A+AS/145/275 Nitta A.: J/A+A/444/643 J/AJ/123/3409 J/AJ/124/1651 J/AJ/126/1023 J/AJ/131/1674 J/AJ/132/676

J/ApJ/607/426 J/ApJS/165/1 J/PASJ/55/771 Nizhel’skii N.A.: J/AZh/84/227 Nizhel’skij N.A.: J/AZh/62/229 Nizhelsky N.A.: J/A+AS/139/545 Nobili S.: J/A+A/437/789 Noble C.J.: J/A+A/446/1185 Noble J.: J/ApJS/115/185 NOBLE R.G.: J/MNRAS/250/314 Noda S.: J/MNRAS/330/137 J/MNRAS/331/609 Noel F.: J/A+A/375/614 NOEL F.: J/A+AS/102/11 J/A+AS/106/327 Noel F.: J/A+AS/106/441 J/A+AS/113/131 J/A+AS/132/195 Noels A.: J/A+A/415/241 NOERGAARD-NIELSEN H.U.: J/A+AS/104/179 Noeske K.G.: J/A+A/429/115 Noguchi C.: J/MNRAS/330/137 Noguchi M.: J/other/PNAOJ/2.561 Noguchi T.: II/106 II/106 Nohejl K.: J/MNRAS/335/1085 Nolan P.L.: J/ApJ/626/95 J/ApJS/101/259 J/ApJS/107/227 J/ApJS/123/79 J/ApJS/94/551 Noll S.: J/A+A/398/49 J/A+A/418/885 J/A+A/420/97 J/A+A/421/41 Nolta M.R.: J/ApJS/148/97 Nolthenious R.: J/AJ/114/1365 Nonino M.: J/A+A/432/381 J/A+A/434/53 J/A+A/444/685 J/A+A/449/951 J/A+A/451/457 J/A+A/454/423

J/A+A/478/83 J/A+AS/137/75 J/A+AS/137/83 J/AJ/105/393 J/ApJ/600/L99 J/ApJ/613/200 J/ApJS/139/369 J/ApJS/155/271 J/ApJS/161/21

Noordermeer E.: J/A+A/442/137 Norberg P.: J/MNRAS/329/227 J/MNRAS/329/87 J/MNRAS/350/1485 J/MNRAS/352/939 Norbury M.A.: VIII/72 Norci L.: J/A+A/458/225 Nord M.E.: J/AJ/128/1646 J/AJ/132/242 Nordh L.: J/A+A/350/883 J/A+A/357/219 J/A+A/372/173 J/A+A/420/945 J/A+A/421/623 Nordsieck K.H.: VI/105 NORDSTORM B.: III/33 NORDSTROEM B.: III/127 III/127 J/A+A/291/121 Nordstroem B.: J/A+A/386/187 J/A+AS/118/407 J/A+AS/126/21 Nordstroem F.: J/A+A/403/1105 NORDSTROM B.: III/102 III/103 III/104 III/105 III/134 III/34 Nordstrom B.: III/35 NORDSTROM B.: J/A+A/299/53 Nordstrom B.: J/A+A/322/460 J/A+A/416/1117 J/AJ/114/1087 V/117 Norgaard-Nielsen H.: J/MNRAS/351/1290 NORGAARD-NIELSEN H.U.: J/A+A/279/61 Noriega-Crespo A.: J/AJ/131/250 Norman C.: J/A+A/451/457 J/ApJ/595/685 J/ApJS/139/369 J/ApJS/155/271 J/ApJS/155/73 Norman C.A.: J/AJ/127/213 Norrington P.H.: J/A+A/416/383 J/A+A/424/363 J/A+A/433/745 J/A+A/455/1157 J/A+A/475/765 Norris J.: J/AJ/125/2502 J/ApJS/61/667 Norris J.E.: J/AJ/117/981 J/ApJ/462/241 J/ApJS/123/639 J/ApJS/93/161 J/ApJS/99/637 J/MNRAS/352/285 Norris N.: J/AJ/120/2496 Norris R.P.: J/A+A/403/1095 J/A+AS/90/327 J/AJ/130/1358 J/AJ/130/1373 J/AJ/131/1216 J/MNRAS/301/1019

J/other/PASAu/10.310 Norsen T.: J/ApJS/162/346 NORTH P.: II/109

North P.: II/58 J/A+A/314/491 J/A+A/319/867 J/A+A/324/137 J/A+A/324/949 J/A+A/326/722 J/A+A/336/953 J/A+A/347/164 J/A+A/360/509 J/A+A/374/204 J/A+A/394/151 J/A+A/398/1121 J/A+A/419/533 J/A+AS/111/41 J/A+AS/122/51 J/A+AS/123/93 J/A+AS/127/277 J/A+AS/130/223 J/A+AS/132/367

Norton A.J.: J/A+A/441/685 J/A+A/467/785 J/MNRAS/372/1117 J/MNRAS/379/647 J/MNRAS/379/816 Noskova R.I.: J/PAZh/25/30 J/PAZh/26/705 J/PAZh/27/187 J/PAZh/27/841 J/PAZh/29/858 J/PAZh/30/297

J/PAZh/30/855 J/PAZh/33/679 VII/236 Nota A.: J/A+A/287/769 J/A+A/288/413 Notni P.: J/AN/322/73 J/AN/325/307 NOTTALE L.: VII/88 Nousek J.: J/A+A/448/L9 Nousek J.A.: J/AJ/122/2810 J/ApJ/554/742 Novak R.: J/A+A/448/1107 Novati S.C.: J/MNRAS/351/1071 Novikov V.V.: III/201 J/PAZh/32/628 J/PAZh/32/672 Novopashenny V.B.: III/201 Nowak M.A.: J/A+A/407/1039 J/A+A/447/245 Nowotny W.: J/A+A/403/93 J/A+A/427/613 Noyelles B.: J/A+A/410/343 Noyes R.W.: J/A+A/410/1051 J/AJ/128/1761 J/PASP/96/707 Noyola E.: J/AJ/132/447 Nozawa S.: J/A+A/417/827 Nucciarelli G.: J/A+A/395/11 J/A+A/402/151 Nugent P.E.: J/A+A/437/789 Nummelin A.: J/A+A/473/177 J/ApJS/117/427 Nummila S.: J/AJ/133/1947 Nunez I.: J/ApJS/126/1 Nunez J.: J/other/NewA/12.461 Nunez J.A.: J/A+A/430/471 Nuritdinov S.N.: J/A+A/420/447 J/A+A/440/53 Nurmi P.: J/A+A/427/769 J/A+AS/132/305 J/A+AS/143/357 Nurnberger D.E.A.: J/ApJ/602/850 Nuspl J.: J/A+A/336/920 J/MNRAS/372/1657 Nutzman P.A.: J/ApJ/626/L5 Nygren D.R.: J/other/APh/26.282

Nyman L.-A.: J/A+A/276/25 J/A+A/376/1123 J/A+A/426/119 J/A+A/426/97 J/A+AS/115/81

J/A+AS/127/185 NYMAN L.A.: J/A+AS/93/121 Nyman L.-A.: VIII/39 Nys O.: I/196 I/211 I/260 I/274

Oasa Y.: J/AJ/131/1608 J/PASJ/58/L29 Oates A.P.: J/MNRAS/300/417 Obayashi H.: J/PASJ/56/295 Oberlack U.: J/A+AS/143/145 OBLAK E.: I/74 Oblak E.: J/A+A/367/865 J/A+A/374/132 J/A+A/446/785 J/A+AS/112/291 J/A+AS/137/451 O’Brien J.: J/AJ/124/1954 J/AJ/128/16 O’Brien J.C.: J/AJ/124/690 O’Brien P.T.: J/A+A/369/57 J/ApJS/115/185 Ocampo A.: I/153 Ocampo H.: I/76 OCHSENBEIN F.: IV/10 IV/11 Ochsenbein F.: IV/6 IV/9 J/ApJ/616/1284 V/100 V/100 V/127 V/15 OCHSENBEIN F.: V/26 Ochsenbein F.: V/53A V/53A V/84 OCHSENBEIN F.: VI/14 Ochsenbein F.: VII/24 O’Connell R.D.: J/AJ/119/1268 O’Connell R.W.: J/AJ/108/1350 J/AJ/108/1598 OConnell R.W.: J/AJ/112/601 O’Connell R.W.: J/AJ/113/1011 J/AJ/116/180 J/AJ/121/768 J/AJ/131/866 J/ApJ/405/94 O’CONNELL R.W.: J/ApJ/413/604 J/ApJ/425/122 O’Connell R.W.: J/ApJ/438/181 OConnell R.W.: J/ApJ/446/622 O’Connell R.W.: J/ApJS/104/287 J/ApJS/82/197 OConnell R.W.: J/ApJS/98/595 O’Connell R.W.: VI/65 O’dea C.P.: J/A+A/463/97 O’Dea C.P.: J/A+AS/131/303 O’dea C.P.: J/AJ/125/478 Odegard N.: J/ApJS/148/97 Odell A.P.: J/A+A/337/183 J/AJ/124/3061 J/ApJS/95/163 O’Dell C.R.: J/AJ/111/846 O’dell C.R.: J/AJ/124/445 J/AJ/125/277 O’Dell C.R.: J/AJ/127/3456 O’Dell M.A.: J/A+AS/121/213 O’Dell S.L.: J/ApJ/615/897 O’DELL S.L.: VIII/26 Odenkirchen M.: J/A+A/400/877 J/A+A/447/173 J/ApJ/588/824 Odenwald S.: J/ApJ/500/554 Odewahn S. C.: J/A+A/470/505 Odewahn S.: J/AJ/107/2120 J/AJ/112/2408 Odewahn S.C.: II/254 J/AJ/125/1660 J/AJ/125/2064 J/AJ/128/1017 J/ApJ/630/784 ODEWAHN S.C.: J/PASP/104/553 Odewahn S.C.: VII/204 Odewhan S.C.: J/A+A/472/121 O’Donoghue A.A.: J/AJ/116/2805 J/AJ/132/161 O’Donoghue D.: J/MNRAS/274/1194 J/MNRAS/287/867 J/MNRAS/320/451 O’dwyer I.J.: J/AJ/125/2239 OEgelman H.: J/other/APh/26.282 OEgelman H.B.: IX/9 OEgerle W.: J/ApJ/416/36 J/ApJ/579/93 OEGERLE W.R.: J/AJ/104/2078 J/AJ/106/831 OEgerle W.R.: J/AJ/116/1529 J/AJ/131/2426 J/ApJ/613/841 J/ApJS/126/133 J/ApJS/146/1 OEmler A.: J/AJ/88/1285 J/ApJ/591/53 VII/203 OEmler A.Jr.: J/AJ/83/1549 OEmler A.Jr: J/ApJ/314/493 OEmler A.Jr.: J/ApJS/110/213 OEmler A.Jr: J/ApJS/122/51 J/ApJS/130/237 OEstlin G.: J/A+A/335/85 J/A+AS/137/419 OEstreicher M.O.: J/A+AS/121/247 J/A+AS/99/591 J/AN/320/105 J/MNRAS/286/500 J/MNRAS/289/729

J/MNRAS/306/279 OEy M.S.: J/AJ/127/1632 J/AJ/129/393 J/ApJ/452/210 J/ApJS/104/71 J/ApJS/165/307 J/ApJS/92/53

J/ApJS/95/1 OEzel M.E.: IX/9 OEzeren F.F.: J/AN/323/31 Ofek E.O.: J/A+A/402/1157 J/MNRAS/287/293 J/MNRAS/382/412 Ogasaka Y.: J/ApJ/518/656 Ogawa H.: J/ApJS/110/21 J/PASJ/51/791 J/PASJ/53/971 Ogloza W.: J/A+A/368/932 J/AJ/130/767 Ognjanovic M.: J/other/SerAJ/166.13Short Ogura: J/A+A/374/504 OGURA K.: III/177 III/177 III/177 III/177 Ogura K.: J/AJ/123/2597 J/MNRAS/358/1290 J/other/BASI/35.53 Oguri M.: J/AJ/131/1 J/MNRAS/382/412 Oh S.J.: J/AJ/126/2840 Ohashi T.: J/ApJS/133/1 J/ApJS/161/185 J/PASJ/53/445 J/PASJ/56/965 O’Hely E.: J/MNRAS/327/588 O’hely E.: J/MNRAS/366/645 Ohishi M.: J/ApJS/117/427 Ohkubo M.: J/PASJ/57/13 J/PASJ/57/27 Ohnaka K.: J/A+A/310/933 J/A+A/353/528 J/A+A/466/1099 J/A+A/470/191 J/A+A/470/957 Ohnishi K.: J/MNRAS/330/137

Ohta K.: J/AJ/111/42 J/ApJ/518/656 J/ApJ/528/123 J/ApJ/610/128 J/ApJ/648/7 J/ApJS/148/275

J/PASJ/53/653 J/PASJ/57/165 J/PASJ/58/313 OHTANI H.: VII/140 VII/140 Ohtsubo J.: J/other/PNAOJ/2.561 Ohyama Y.: J/PASJ/57/165 OJA T.: II/182 Oja T.: II/21A J/A+AS/113/115 J/A+AS/139/513 Ojala T.: J/AJ/133/1947

Ojha D.K.: J/A+A/403/141 J/A+A/405/531 J/A+A/463/175 J/A+A/468/1001 J/MNRAS/381/1219

J/other/BASI/35.53 Ojha R.: J/AJ/127/3609 J/AJ/128/1570 J/AJ/130/2529 J/ApJ/614/607 Oka T.: J/ApJS/100/125 Okamura S.: J/AJ/121/2331 J/AJ/125/1817 J/AJ/127/180 J/ApJ/648/7 J/ApJS/137/279 J/ApJS/97/77

J/MNRAS/316/326 J/MNRAS/357/1357 J/PASJ/55/1079 J/PASJ/55/771 J/PASJ/57/165 J/PASJ/58/313 VII/147

Okazaki A.: J/A+A/327/145 Oke J.B.: J/AJ/116/560 J/AJ/116/584 J/AJ/122/1125 J/ApJ/306/30 J/ApJ/471/694 Okhmat N.N.: J/other/Ap/48.304 Oksanen J.: J/A+AS/132/305 Okubo H.: J/PASJ/52/631 Okuda H.: J/A+A/365/409 J/A+A/413/843 J/A+A/449/583 Okudaira A.: J/PASJ/52/631 Okumura K.: J/A+A/407/791 OLAH K.: J/A+A/291/110 Olbrechts PH.: J/other/APh/26.282 Olech A.: J/A+A/424/1101 J/AcA/46/389 Oliva E.: J/AJ/124/3222 Olivares J.I.: J/A+AS/129/9 Olivas A.: J/other/APh/26.282 Olive J.-F.: J/other/GCN/1070.1 Oliveira C.M.: J/A+A/470/317 Oliveira H.L.: J/AJ/117/238 Oliveira: J/A+A/444/895 Oliveira J.M.: J/MNRAS/356/89 J/MNRAS/358/13 J/MNRAS/376/580 OLIVER J.P.: VI/23 Oliver J.P.: VI/44 Oliver R.J.: J/A+A/315/578 Oliver S.: J/A+A/372/364 J/A+A/451/431 J/AJ/127/3075 J/AJ/129/1198 J/AJ/129/2074 J/AJ/131/2383

J/MNRAS/302/222 J/MNRAS/305/297 J/MNRAS/325/1173 J/MNRAS/344/887 J/MNRAS/351/1290 J/MNRAS/355/97 J/MNRAS/358/333 J/MNRAS/358/397 J/MNRAS/372/1621

Oliver S.J.: J/MNRAS/343/293 VII/221 Olivier E.A.: J/MNRAS/376/313 Olmi L.: J/A+AS/103/541

Olofsson G.: J/A+A/350/883 J/A+A/357/219 J/A+A/372/173 J/A+A/376/907 J/A+A/399/995

J/A+A/420/945 J/A+A/421/623 J/A+A/458/789 J/A+A/475/281 Olofsson H.: J/A+A/403/93 J/A+A/427/613 J/A+AS/127/185 Olofsson K.: J/A+A/321/29 OLOFSSON K.: J/A+AS/111/57 Olofsson S.: J/A+A/420/945 Olowin R.: J/ApJ/520/491 Olowin R.P.: VII/110A OLSEN E.H.: II/100 II/138 Olsen E.H.: II/30 II/33 II/78 II/90 J/A+A/352/600 J/A+A/352/605 J/A+A/392/215 J/A+A/461/1065 OLSEN E.H.: J/A+AS/102/89 J/A+AS/104/429 J/A+AS/106/257 Olsen E.H.: J/A+AS/145/275 J/A+AS/146/365 V/117 Olsen H.J. Fogh: I/99 I/99 I/99 OLSEN H.J.F.: I/14 I/23 I/30 Olsen K.: J/ApJ/597/323 Olsen K.A.G.: J/AJ/117/2244 J/AJ/126/2867 J/AJ/127/2674 J/AJ/131/1416 J/AJ/131/2478 J/AJ/132/271

J/MNRAS/300/665 Olsen L.F.: J/A+A/384/81 J/A+A/388/1 J/A+A/392/741 J/A+A/409/439 J/A+A/435/781 J/A+A/442/841

J/A+A/452/119 J/A+A/456/881 J/A+A/461/81 J/A+A/462/553 J/A+AS/137/75 J/A+AS/137/83 Olson E.C.: J/A+AS/132/367 J/AJ/109/1269 J/AJ/109/1308 J/ApJ/484/394 Olszewski E.: J/AJ/120/2496 Olszewski E.W.: J/AJ/110/1186 J/AJ/110/2131 J/AJ/110/652 J/AJ/121/841 J/AJ/123/2511 J/AJ/125/2502

J/AJ/131/2114 J/AJ/131/912 J/AJ/90/2221 J/ApJ/562/303 J/MNRAS/300/665 O’Mara B.J.: J/A+AS/142/467 Omizzolo A.: J/AJ/119/2540 J/AJ/124/2955 VII/69 Omodaka T.: J/PASJ/52/631

Omont A.: II/243 J/A+A/267/515 J/A+A/275/163 J/A+A/285/247 J/A+A/319/235 J/A+A/349/L69 J/A+A/357/225 J/A+A/362/199 J/A+A/362/215 J/A+A/392/971 J/A+A/393/115 J/A+A/400/903 J/A+A/403/141 J/A+A/405/531 J/A+A/406/51 J/A+A/418/103 J/A+A/435/575 J/A+A/446/551 J/A+A/453/535 J/A+AS/99/291 J/ApJ/517/148 J/ApJ/552/289 J/MNRAS/351/1290 J/MNRAS/381/1219

Ona E.: J/A+A/390/39 Onaka T.: J/A+A/423/579 J/ApJS/155/401 Ona-Wilhelmi E.: J/A+A/406/L9 J/A+A/410/813 O’Neal R.H.JR: J/ApJS/157/324 O’Neil E.J.: J/AJ/113/1691 J/ApJ/478/124 O’Neil E.J.Jr: J/AJ/116/146 O’Neil E.J.JR.: J/ApJ/456/174 J/ApJ/459/L27 O’Neil E.J.Jr: J/ApJ/468/633 Oneil Jr.: J/ApJ/448/179 O’Neil K.: J/A+A/408/67 O’neil K.: J/A+A/449/929 O’Neil K.: J/AJ/128/2080 J/ApJ/613/914 J/ApJS/128/99 O’neil K.: J/MNRAS/352/478 J/MNRAS/357/819 O’Neil K.O.: J/A+A/428/823 Onishi T.: J/PASJ/53/1081 J/PASJ/53/971 J/PASJ/57/917 Onken C.: J/AJ/121/318 Onken C.A.: J/ApJ/613/682 Oort M.J.A.: J/A+AS/71/221 J/A+AS/73/103 Oosterbroeck T.: J/A+A/242/401 Oosterloo T.: J/AJ/123/873 J/AJ/124/1954 J/AJ/124/690 J/AJ/128/16 Oosterloo T.A.: J/A+A/453/493 Oppenheimer B.D.: J/AJ/129/2294 Oppenheimer B.R.: IX/18 J/AJ/121/2189 J/other/Sci/292.698 Oprescu G.: J/A+A/327/1017 J/A+AS/125/399 Ordonez R.: J/other/RMxAA/35.187 Orellana R.B.: J/other/NewA/12.461 Orfei A.: J/A+A/379/755 Orienti M.: J/A+A/426/463 J/A+A/450/959

Origlia L.: J/A+A/298/461 J/A+A/419/139 J/A+A/420/173 J/AJ/124/3222 J/MNRAS/351/1204

J/MNRAS/361/272 Origlia M.: J/ApJ/654/915 Orio M.: J/A+A/389/439 Orlov V.G.: J/A+AS/105/503 Orlov V.V.: J/AJ/117/587 J/AZh/72/495 J/AZh/82/231 J/PAZh/26/301 Orlova L.F.: J/AZh/84/429 Orofino V.: J/A+A/330/505 Orosz J.A.: J/AJ/113/1045 Orsatti A.M.: J/A+A/471/165 J/A+AS/144/195 J/AJ/82/345 J/AJ/82/474 Ortiz J.: J/A+A/407/1139 Ortiz J.L.: J/A+A/468/L13 Ortiz P.: J/AJ/112/2408 Ortiz P.F.: J/AJ/124/690 ORTIZ R.: J/A+A/299/453 Ortiz R.: J/A+A/362/215 J/A+A/431/565 J/A+A/461/949 Ortiz-Gil A.: J/A+A/425/367 ORTOLANI S.: J/A+A/286/444 J/A+A/296/680 J/A+A/300/128 J/A+A/300/726 Ortolani S.: J/A+A/306/134 J/A+A/308/733 J/A+A/319/850 J/A+A/326/614 J/A+A/331/70 J/A+A/372/879

J/A+A/390/931 J/A+A/391/179 J/A+A/401/997 J/A+A/429/607 J/A+A/435/657 J/A+A/444/831 J/A+A/450/105 J/A+A/464/953 J/A+A/465/799 J/A+A/467/1025 J/A+AS/106/161

ORTOLANI S.: J/A+AS/106/573 J/A+AS/108/571 J/A+AS/108/653 J/A+AS/111/17

Ortolani S.: J/A+AS/120/153 J/A+AS/122/483 J/A+AS/126/319 J/A+AS/127/471 J/A+AS/145/399

J/A+AS/82/71 J/MNRAS/289/745 Osaki H.: J/PASJ/52/631 Osamura Y.: J/ApJ/519/697 Osborn W.: J/A+AS/24/35 J/AJ/89/1897 J/other/Obs/116.382 J/PASP/114/1382 Osborn W.H.: VII/11 Osborne J.: J/MNRAS/372/1117 J/MNRAS/379/647 J/MNRAS/379/816 Osborne J.L.: J/MNRAS/201/495 Osborne J.P.: J/A+A/448/L9 J/A+A/476/1191 Osborne L.: VI/43 Osip D.: J/ApJ/626/L5

Osmer P.S.: J/AJ/114/2269 J/AJ/121/1840 J/ApJ/462/614 J/ApJS/104/185 J/ApJS/119/189 VII/148

VII/182 VII/182 Ossenkopf V.: J/A+A/291/943 Ostensen R.: J/A+A/459/557 Ostensen R.H.: J/A+A/450/1149 Osterbrock D.E.: III/211 Osterloh M.: J/A+AS/120/267 Ostheimer J.C.: J/AJ/125/1352 J/AJ/130/2677 Ostorero L.: J/A+A/395/11 J/A+A/402/151 J/A+AS/144/481 OSTROV P.: J/AJ/105/1762 Ostrov P.G.: J/A+A/380/258 O’Sullivan E.: J/MNRAS/328/461 O’sullivan E.: J/MNRAS/367/627 O’SULLIVAN S.S.J.: II/194A Oswalt T.D.: J/A+A/459/557 J/AJ/124/1118 J/AJ/129/2428 J/ApJ/571/512 J/ApJS/115/185 Otero S.A.: J/other/IBVS/5480 J/other/IBVS/5495 J/other/IBVS/5586 Otruba W.: J/A+A/447/735 OTRUPCEK R.: J/MNRAS/251/330 Otrupcek R.: VIII/15 Otrupcek R.E.: J/other/PASA/17.92 Ott S.: J/A+A/407/791 Ott T.: J/ApJ/523/248 J/ApJ/643/1011 Ottaviani D.L.: J/ApJS/111/377 Ouchi M.: J/ApJ/634/861 J/ApJ/648/7 J/MNRAS/357/1357 J/PASJ/55/1079 J/PASJ/57/165 J/PASJ/58/313 Oudmaijer R.: J/A+A/365/110

Oudmaijer R.D.: J/A+A/378/116 J/A+A/379/564 J/A+A/439/1107 J/A+A/461/11 J/A+A/474/891

J/A+AS/129/541 J/MNRAS/363/1111 J/MNRAS/366/1096 Oukbir J.: J/A+AS/124/283 Ounnas C.: J/A+AS/49/715 Ounnas CH.: J/A+AS/76/205 Outram P.: J/MNRAS/372/425 Outram P.J.: J/MNRAS/328/805 Ovcharov E.P.: J/A+A/466/501 Ovchinnikov S.L.: III/208 Overgaard L.P.: J/A+AS/117/317 Overzier R.: J/AJ/128/1990 Overzier R.A.: J/A+A/431/793 J/A+A/461/823 Owen F.: J/AJ/129/2074 J/ApJ/611/821

Owen F.N.: J/AJ/109/14 J/AJ/109/853 J/AJ/110/1959 J/AJ/118/625 J/AJ/123/3018 J/AJ/124/2453 J/AJ/125/2393 J/AJ/125/2427 J/AJ/125/478 J/AJ/126/119 J/AJ/126/2740 J/AJ/129/26 J/AJ/129/31 J/AJ/130/47 J/AJ/98/64 J/ApJ/581/844 J/ApJ/613/841 J/ApJS/134/355 J/ApJS/146/267 J/ApJS/80/501

OWEN F.N: VIII/29 OWEN Jr W.M.: I/140 Owen R.: J/AJ/121/2308 J/AJ/121/2331 J/AJ/123/567 Oyabu S.: J/A+A/365/409 J/A+A/413/843 Oyama T.: J/PASJ/52/631 Ozawa H.: J/A+A/429/963 Ozdemir S.: J/A+A/387/240 J/A+A/403/675 J/PAZh/26/843 Ozel F.: J/ApJ/615/887 OZOLINA V.: III/140

P.: J/A+A/384/371 J/A+A/428/587 J/A+A/428/587 J/A+AS/136/509 J/A+AS/136/509 J/A+AS/139/483

J/A+AS/139/483 J/MNRAS/273/1097 Paardekooper S.J.: J/A+A/384/1012 J/A+A/404/L29 Pacaud F.: J/A+A/474/473 J/MNRAS/382/279 Pace G.: J/A+A/401/997 Pacheco A.M.: J/A+A/430/327 J/A+AS/136/1 Pacheco R.: J/A+A/465/331 Paciesas W.S.: IX/20A J/ApJS/126/19 J/ApJS/154/585 Packham C.: J/MNRAS/323/839 J/MNRAS/330/1 J/MNRAS/346/1125 Paczynski B.: II/278 J/AcA/49/319 J/ApJ/636/240 J/MNRAS/349/1059 Padgett D.: J/AJ/129/2074 J/PASP/118/1690 Padgett D.L.: J/A+A/468/353 J/AJ/131/250 Padilla L.: J/A+A/390/39 Padoan F.: J/A+A/433/497

Padovani P.: J/A+A/424/545 J/A+A/434/385 J/A+A/470/787 J/AJ/115/1253 J/MNRAS/277/1477

J/MNRAS/323/757 J/MNRAS/351/83 Padrielli L.: J/A+AS/120/89 Pagani C.: J/A+A/448/L9 Pagani L.: J/A+A/328/617 Pagano I.: J/A+A/310/173 J/A+A/327/1114 J/A+A/415/331 J/A+A/469/713 PAGE A.A.: III/110 III/71 Page C.G.: J/MNRAS/274/1165 VII/19 Page K.L.: J/A+A/448/L9 J/A+A/474/793 Page L.: J/ApJS/148/97 Page M.J.: J/A+A/428/383 J/A+A/469/27 J/A+A/474/473 J/A+A/476/1191 J/ApJS/154/155 J/MNRAS/311/456

J/MNRAS/318/1073 J/MNRAS/342/802 J/MNRAS/362/1371 J/MNRAS/369/156 J/MNRAS/372/1621 J/MNRAS/378/995

Page T.: II/92 Pahre M. A.: J/ApJS/124/127 Pahre M.A.: J/AJ/131/527 J/ApJ/560/566 J/ApJS/127/1 PAIK Y.: J/AJ/107/1022 Pain R.: J/A+A/437/789 J/ApJ/645/488 Paioro L.: J/A+A/439/413 J/MNRAS/382/279 Paizis A.: J/ApJ/635/214 PAJDOSZ G.: J/A+A/285/459 Pak Soojong: J/A+A/450/39 Pakhomov V.P.: III/201 J/PAZh/32/672 Pakhomov Yu.V.: J/AZh/79/909 J/AZh/80/704 J/AZh/81/658 J/AZh/82/601 Pakuliak L.: I/261 Pakull M.: J/ApJS/117/319 Pakull M.W.: J/A+A/318/111 J/A+AS/122/201 J/A+AS/132/341 PAKULYAK L.K.: I/192 Pakulyak L.K.: J/AZh/80/291 Palacios J.: J/A+A/365/110 J/A+A/378/116 J/A+A/379/564 Paladini R.: J/A+A/397/213 PALAGI F.: J/A+A/291/261 Palagi F.: J/A+A/368/845 J/A+AS/101/153 J/A+AS/103/541 J/A+AS/134/115 Palanque-Delabrouille N.: J/A+A/389/149 J/A+A/454/185 Palau A.: J/A+A/474/911 Palazzi E.: J/A+A/469/503 J/A+A/474/793 Paldino R.: J/A+A/375/130 Paletou F.: J/A+A/475/1053 PALLA F.: J/A+A/291/261 Palla F.: J/A+A/308/573 J/A+A/327/755 J/A+A/336/339 J/A+A/368/845 J/A+A/399/271 J/A+A/448/189

J/A+A/450/147 J/A+A/468/353 J/A+AS/103/541 J/ApJ/581/1194 Pallavicini R.: J/A+A/326/249 J/A+A/372/862 J/A+A/446/501 J/A+AS/115/41 J/A+AS/138/87 Palle P.L.: J/A+A/390/717 Palma C.: J/AJ/125/1352 J/AJ/130/2677 Palmeri P.: J/A+A/410/359 J/A+A/414/377 J/A+A/418/1171 Palomar Obs.: VI/25 PALOQUE E.: I/55 Palous J.: J/A+A/437/101 V/30 Paltani S.: J/A+A/428/1043 J/A+A/457/79 Paltrinieri B.: J/A+A/320/757 J/A+A/324/915 J/AJ/121/3114 J/ApJ/522/983 Palumbo G.: J/A+AS/64/247 Palumbo G.G.C: J/A+A/335/912 Palumbo G.G.C.: J/A+A/336/339 J/ApJS/106/399 J/MNRAS/276/57 PALUMBO G.G.C.: VII/181 Palumbo G.G.C.: VII/213 PALUMBO G.G.C.: VII/70A Palumbo G.G.C.: VII/71 VII/99 Palunas P.: J/ApJ/548/127 J/ApJ/614/75 Palviainen A.: J/A+A/383/197 Pan J.: J/ApJ/499/517 Pan K.: J/ApJS/151/313 Pan M.: J/A+A/465/331 Pan R.S.: J/A+AS/100/243 Panagi P.M.: J/A+AS/121/213 PANAGIA N.: III/167 Panagia N.: J/A+A/397/473 J/AJ/129/1534 J/ApJ/522/802 J/ApJ/551/973 J/ApJ/613/200 J/ApJS/99/223 Panchenko I.E.: J/AZh/72/495

Panchuk V.E.: III/243 J/A+A/381/32 J/A+A/397/1035 J/A+A/425/671 J/AZh/80/704 J/AZh/81/658

J/AZh/83/265 J/PAZh/26/115 J/PAZh/26/465 J/PAZh/26/510 Pancino E.: J/A+A/420/173 J/A+A/455/271 J/A+A/464/953 J/AJ/124/3222 J/ApJ/654/915 J/MNRAS/349/1278

J/MNRAS/357/265 Pandey A.K.: J/A+A/372/71 J/A+A/402/113 J/A+AS/122/111 J/MNRAS/358/1290 J/MNRAS/379/1237

J/other/BASI/35.53 Pandey G.: J/MNRAS/369/1677 Pandey: J/A+A/374/504 Panek R.J.: III/115 Panella D.: J/A+A/368/845 J/A+AS/103/541 Paneque D.: J/A+A/410/813 Panessa F.: J/A+A/476/1191 Panis J.F.: J/A+AS/122/271 Panis J.-F.: VIII/66 VIII/66 PANJAITAN E.: I/128 Pannella M.: J/A+A/428/339 J/MNRAS/372/1425 Pannuti T.G.: J/A+A/435/437 J/ApJS/144/213 Panov K.: J/A+A/467/229 PANOV K.P.: J/A+AS/103/183 Pansecchi L.: J/MNRAS/367/1170 Panter M.: J/A+A/390/39 J/A+A/406/L9 J/A+A/410/813 Pantin E.: J/A+A/473/369 Pantoja C.A.: J/AJ/113/905 Panzera M.R.: IX/34 Papadaki C.: J/A+A/456/599 Papadakis I.E.: J/A+A/402/151 Papaderos P.: J/A+A/421/539 J/A+A/429/115 J/A+A/464/885 J/AJ/121/169 J/ApJ/510/659 Papadopoulos P.: J/MNRAS/351/1290 Papaj J.: II/175 Papamastorakis J.: J/A+A/402/1157 Paparo M.: J/A+A/336/920 J/A+AS/129/9 Papovich C.: J/AJ/126/1183 J/AJ/132/231 J/ApJ/600/L99 J/ApJ/640/167 J/ApJS/154/155 Paraficz D.: J/A+A/455/L1 Paragi Z.: J/ApJS/131/95 J/ApJS/155/33 J/PASJ/52/997 Pardo J.R.: J/A+A/329/219 Paredes J.M.: J/A+A/472/557 Parenago P.P.: II/171 Paresce F.: J/A+A/287/769 J/A+A/288/413 J/A+A/413/1037 J/AJ/106/154 J/AJ/121/3114 PARESCE F.: J/ApJS/85/293 Pareschi G.: J/A+AS/103/541 J/MNRAS/276/57 Parihar P.S.: J/A+A/469/713 Pariiskii Yu.N.: J/AZh/73/835 J/AZh/79/589 Parijskij Yu.N.: J/A+AS/87/1 J/AZh/71/684 J/AZh/72/613 J/AZh/84/227 J/other/BSAO/40.5 PARIS: I/37 I/67 Parisi M.C.: J/A+A/440/111 J/MNRAS/377/1737 PARISOT J.P.: I/74 Park B.-G.: J/A+A/459/557 Park H.S.: J/A+A/399/365 J/AJ/126/2840 Park J.-C.: J/other/NewA/11.293 Park J.-H.: J/AJ/126/2840 Park M.-G.: J/PASP/118/636 Park S.Q.: J/ApJ/642/126 Park W.-K.: J/AJ/126/2840 Park Y.: J/ApJ/600/L155 Parker J.E.: J/AJ/127/1567

Parker J.W.: III/188 J/AJ/108/1256 J/AJ/116/180 J/AJ/118/1684 J/AJ/121/891 J/AJ/127/1632

J/AJ/98/1305 PARKER J.W.M.: II/187 Parker J.WM.: J/AJ/103/1205 Parker J.Wm.: J/AJ/113/1011 Parker J.Wn.: J/AJ/111/1128 Parker Q.A.: J/A+A/476/59 J/MNRAS/283/367 J/MNRAS/287/472 J/MNRAS/300/417 J/MNRAS/306/592

J/MNRAS/316/326 J/MNRAS/346/787 J/MNRAS/355/1303 J/MNRAS/362/689 J/MNRAS/365/401 J/MNRAS/373/521 J/other/PASA/16.113 J/other/PASA/18.232 V/127

Parkinson W.H: VI/61 Parley N.: J/A+A/467/785 Parley N.R.: J/MNRAS/379/647 Parma P.: J/A+A/339/34 J/A+A/403/857 J/A+A/439/487 J/A+A/457/517 J/A+AS/106/1 J/A+AS/95/1

VIII/63 Parmar A.N.: J/A+AS/109/9 J/A+AS/134/287 Parmeggiani G.: J/A+A/473/429 Parnovsky S.L.: J/other/BSAO/50.5 J/other/KFNT/22.187 VII/162 VII/219 Parodi B.R.: J/A+A/388/29 PARRA F.: J/A+AS/110/351 Parrao: J/A+A/445/939 Parrao L.: J/A+A/426/247 J/A+A/438/201 J/A+A/468/637 J/A+AS/117/317 J/A+AS/129/9 J/A+AS/97/951 Parry I.R.: J/MNRAS/308/897 Parsons S.B.: III/86 J/AJ/127/2915 J/ApJS/119/83 PARSONS S.B.: VI/15 VI/32 Parthasarathy M.: J/A+A/297/359 J/A+A/385/884 J/A+A/409/251 J/A+A/423/353 PARTHASARATHY M.: J/A+AS/111/407 Partridge R.B.: J/AJ/123/2402 PARTRIDGE R.B.: J/ApJ/405/498 Partridge R.B.: J/ApJ/635/950 J/MNRAS/358/1159 Parviainen M.: J/AJ/133/1947 Pascarelle S.M.: J/ApJ/538/493 J/ApJS/135/41 Paschke A.: J/other/IBVS/5731 Pascu D.: J/AJ/124/2254 J/AJ/127/539 Pascual S.: J/A+A/379/798 Pascucci I.: J/PASP/118/1690 Pasinetti L.E.: II/61 PASINETTI-FRACASSINI L.E.: II/155 Pasinetti-Fracassini L.E.: II/224 Pasquali A.: J/A+A/280/581 J/A+A/391/857

Pasquini L.: J/A+A/319/593 J/A+A/326/249 J/A+A/330/453 J/A+A/361/1011 J/A+A/384/491 J/A+A/384/491 J/A+A/401/997 J/A+A/413/343 J/A+A/414/503 J/A+A/435/657 J/A+A/464/953 J/A+AS/115/41 J/A+AS/131/197 J/MNRAS/327/918

Passuello R.: J/A+AS/116/1 Pastor J.: J/ApJ/463/205 Pastorello A.: J/MNRAS/369/1880 PASTORI L.: II/155 Pastori L.: II/224 J/A+A/364/205 J/A+A/384/491 J/A+A/384/491 J/A+AS/138/87 Pastoriza M.: J/ApJS/127/39 Pastoriza M.G.: J/A+A/428/837 J/A+A/455/453 J/A+A/467/1003 Pastukhova E.N.: II/205 J/other/IBVS/5721 J/PAZh/28/201 Patat F.: J/A+A/317/423 J/A+A/379/136 J/A+A/418/525 J/A+AS/114/281 J/AJ/130/652 J/MNRAS/369/1880 Patchick D.: J/A+A/447/921 Patel K.: J/AJ/110/2715 Patel S.G.: J/AJ/129/1534 Patel S.K.: J/ApJ/563/L49 J/ApJ/615/887 Pati A.: J/A+A/409/251 Pati A.K.: J/A+AS/114/537 Patience J.: J/AJ/123/1570 Patkos L.: J/A+A/417/745 J/AN/325/424 Patnaik A.R.: J/A+AS/135/273 J/ApJ/481/95 J/ApJS/145/83 VIII/60 Patriarchi P.: III/130 J/A+A/410/905 Patsis P.A.: J/A+A/423/849 Patten B.: J/ApJS/134/103 Patten B.M.: J/AJ/119/3026 J/AJ/123/1528 J/AJ/127/2228 J/AJ/131/527 Patterer R.J.: J/AJ/107/751 Patterson J.: J/MNRAS/379/289 Patterson R.J.: J/AJ/125/1352 J/AJ/128/245 J/AJ/130/2677 Patterson R.S.: J/A+A/459/557 Patton D.R.: J/AJ/130/2043 J/ApJ/565/208 J/ApJS/129/475 Patton S.: J/other/APh/26.282 Paturel G.: J/A+A/113/61 J/A+A/363/62 J/A+A/383/398 J/A+A/387/1 J/A+A/393/57 J/A+A/430/373

J/A+AS/102/57 PATUREL G.: J/A+AS/113/151

Paturel G.: J/A+AS/117/467 J/A+AS/130/333 J/A+AS/135/133 J/A+AS/140/89 J/A+AS/144/475 J/A+AS/91/371 J/MNRAS/283/367 J/MNRAS/287/472 VII/119 VII/136 VII/137B VII/155 VII/234 VII/237 VII/238 VII/242

Pau S.: J/A+A/468/637 Paubert G.: J/A+AS/102/611 J/A+AS/122/271 Paul J.F.: J/A+AS/99/291 Pauldrach A.W.A.: J/A+A/321/531 J/A+A/437/467 J/ApJ/566/880 Pauley Brandy R.: J/AJ/113/401 Pauli E.-M.: J/A+A/400/877 J/A+A/447/173 Paulin-henriksson S.: J/MNRAS/351/1071 Pauliny-Toth I.I.K.: J/A+A/85/329 J/A+AS/82/41 J/AJ/77/797 J/ApJS/13/65 PAULINYTOTH I.I.K.: VII/78 VIII/5 Pauls T.: J/A+AS/35/23 Paulson D.B.: J/AJ/127/3579 Paumard T.: J/A+A/426/81 J/ApJ/643/1011

Paunzen E.: J/A+A/331/633 J/A+A/371/L5 J/A+A/373/153 J/A+A/373/625 J/A+A/385/867 J/A+A/395/823 J/A+A/403/937 J/A+A/412/721 J/A+A/441/631 J/A+A/443/157 J/A+A/444/941 J/A+A/454/171 J/A+A/454/179 J/A+A/458/293 J/A+A/459/871 J/A+A/462/157 J/A+A/462/591 J/A+AS/123/93 J/A+AS/128/573 J/A+AS/133/1 J/MNRAS/362/1025

PAUPERS O.: II/198 Paupers O.: III/227 Paust N.E.Q.: J/AJ/131/2538 Pauwels T.: J/A+A/437/705 Pauzat F.: J/A+A/319/318 Pavlenko E.P.: J/AZh/75/231 Pavlenko Y.V: J/A+A/416/655 Pavlenko Y.V.: VI/121 Pavlov G.G.: J/ApJ/578/405 Pavlov M.: J/A+A/428/339 J/A+AS/109/479 J/A+AS/109/487 Pavlov M.V.: J/PAZh/22/269 J/PAZh/22/764 Pavlovic R.: J/A+AS/123/487 PAVLOVSKAYA E.D.: I/150 Pavlovski: J/A+A/454/855 Pavlovski K.: J/A+A/324/965 J/A+A/358/553 PAVLOVSKI K.: J/A+AS/107/403 Pavlovski K.: J/A+AS/125/75 Pavlovsky C.M.: J/AJ/119/3026 Pavlyuk N.N.: II/201 II/218 II/256 J/PAZh/30/37 Pawl A.: J/AJ/119/1901 Paxton L.: II/269 PAYNE H.E.: VIII/26 Payne J.L.: J/MNRAS/355/44 J/other/SerAJ/166.13Short Pazderski E.: J/A+A/474/1093 P.C.: J/MNRAS/325/1002 Peach J.V.: J/MNRAS/239/459 J/MNRAS/240/785 J/MNRAS/267/431 VII/42A Peacock A.: J/A+AS/109/9 J/A+AS/134/287

Peacock J.A.: II/211 J/MNRAS/216/173 J/MNRAS/238/1171 J/MNRAS/238/709 J/MNRAS/317/801 J/MNRAS/329/227 J/MNRAS/329/87 J/MNRAS/333/100 J/MNRAS/344/887 J/MNRAS/350/1485 J/MNRAS/352/939 J/MNRAS/372/1621

Pearson C.P.: J/MNRAS/372/1621 Pearson J.C.: J/ApJS/162/428 Pearson T.J.: J/A+A/472/763 J/ApJS/100/1 J/ApJS/95/345 J/ApJS/98/1 J/ApJS/98/33 J/ApJS/99/297

J/MNRAS/171/475 J/MNRAS/182/273 J/MNRAS/354/485 VIII/72 Pease G.: V/31A Peat D.W.: II/22 II/24 II/25 PEAT D.W.: III/25 III/26 Pebley S.: J/A+A/371/79 Pebley S.P.: J/ApJS/115/185 Pecci F.F.: J/AJ/122/2569 Peck A.B.: J/ApJS/141/13 J/PASJ/57/135 Pedani M.: J/A+A/468/807 J/A+AS/123/219 J/MNRAS/348/857 J/MNRAS/370/1034 Pedotti P.: J/A+A/396/449 J/A+A/446/839 PEDOUSSAUT A.: IV/16 V/49 Pedoussaut A.: V/60 PEDRANA M.D.: J/A+AS/103/245 Pedraz S.: J/A+AS/139/29 J/MNRAS/376/125 Pedretti E.: J/ApJ/620/961 Pegourie B.: J/A+A/293/833 Pei C.C.: J/A+AS/127/181 J/other/PPMtO/14.185 Peila A.: J/A+AS/121/119 J/A+AS/127/445 J/A+AS/129/577 J/A+AS/130/109 Peimbert A.: J/ApJ/595/247 J/MNRAS/362/301 J/MNRAS/368/253 Peimbert M.: III/84B J/ApJ/513/168 J/ApJ/595/247 J/ApJS/120/113 J/MNRAS/362/301 J/MNRAS/368/253 Pekunlu R.: J/A+A/336/920 Peletier: J/MNRAS/326/959 Peletier R.: J/ApJS/111/203 Peletier R.F.: J/A+A/414/23 PELETIER R.F.: J/A+AS/108/621 J/AJ/100/1091 J/AJ/107/135 Peletier R.F.: J/MNRAS/317/234 J/MNRAS/371/703 J/MNRAS/374/664 Pellegrini P.S.: J/AJ/105/1637 J/AJ/108/1987 J/AJ/116/1 J/AJ/125/2307 J/AJ/126/2268 Pellerin A.: J/ApJS/144/21 Pellet A.: J/A+AS/31/439 Pellizza L.: J/A+AS/109/201

Pello R.: III/232 J/A+A/427/397 J/A+A/428/1043 J/A+A/444/365 J/A+A/451/1159 J/A+A/457/79 J/A+A/459/679 J/A+AS/115/219 J/A+AS/141/113 J/A+AS/95/87 J/ApJ/630/206 J/MNRAS/330/1 J/MNRAS/371/1777

PELT J.: J/A+A/278/449 Pelt J.: J/A+A/351/212 J/A+A/394/505 PEMBERTON A.C.: III/25 III/26 Pena J.: J/A+A/468/637 Pena J.H.: J/A+A/426/247 J/A+A/438/201 J/A+AS/129/9 Pena M.: J/A+A/466/75 PENA M.: J/AJ/105/35 Pena M.: J/AJ/108/405 Pena-Garay C.: J/other/APh/26.282 PENCE J.D.: VII/39A Pence W.D.: J/ApJ/561/189 Pendleton G.N.: IX/20A J/ApJS/126/19 J/ApJS/134/385 Penent F.: J/A+A/439/387 Penfold J.E.: J/AJ/110/2400 Peng B.: J/AJ/129/1818 J/other/AcApS/16.103 VIII/44 Peng E.: J/AJ/123/3409 Peng E.W.: J/ApJ/602/685 J/ApJ/613/279 J/ApJS/150/367 J/ApJS/153/223 J/ApJS/154/509 Peng Q.: J/A+A/383/296 J/A+AS/126/503 Peng R.: J/ApJ/604/717 Penhallow W.S.: J/AJ/77/486 Peniche R.: J/A+AS/129/9 Penna J.L.: J/A+A/384/650 J/A+AS/135/227 J/A+AS/136/531 J/A+AS/143/265 Pennington R.L.: J/AJ/96/1884 Penny A.: J/A+A/365/110 J/A+A/378/116 J/A+A/379/564 J/MNRAS/358/795 Penny A.J.: J/A+A/456/651 PENNY A.J.: J/MNRAS/250/314 Penny A.J.: J/MNRAS/370/163 J/MNRAS/372/163 Penprase B.E.: J/A+A/364/712 J/ApJS/83/273 Penston M.J.: I/85 V/32A V/4 Pentericci L.: J/A+A/431/793 J/A+A/461/823 Pepe F.: J/A+A/356/590 J/A+A/375/205 J/A+A/379/999 J/A+A/388/632 J/A+A/390/267 J/A+A/392/215

J/A+A/406/373 J/A+A/415/391 J/A+A/468/1115 J/A+A/474/293 J/A+A/475/1125 PEQUIGNOT D.: J/A+A/303/175 Pequignot D.: J/A+A/442/249 PERAULT M.: II/148

Perault M.: J/A+A/350/883 J/A+A/357/219 J/A+A/372/173 J/A+A/420/945 J/A+A/421/623

J/ApJ/552/289 VIII/66 Percheron I.: J/A+A/386/492 J/A+A/413/1037 J/A+A/431/773 J/A+A/434/1201 Percival J.W.: II/249 Percival S.M.: J/A+A/414/23 Percival W.: J/MNRAS/329/87 Percival W.J.: J/MNRAS/372/1621 Percy J.: J/A+A/387/580 J/A+A/416/669 Percy J.R.: J/A+A/328/551 J/A+A/464/263 Perdereau O.: J/A+A/389/149 J/A+A/454/185 Perdomo R.: J/A+A/430/327 J/A+AS/118/1 J/A+AS/136/1 J/other/PBeiO/24.71 Perea Calderon J.V.: J/A+A/446/773 Perea J.: J/A+A/285/404 J/A+A/310/31 J/A+A/310/8 J/A+AS/129/399 Perea J.V.: J/A+AS/147/5 Pereira C.B.: J/A+A/333/658 J/A+A/477/877 Pereira P.C.R.: J/A+AS/128/561 Perelmuter J.M.: J/AJ/109/1055 Peres G.: J/A+A/430/287 J/A+A/439/1149 J/A+A/456/977 J/A+A/459/199 J/A+A/462/245 Peretto N.: J/A+A/472/519 Perevozkina E.L.: J/A+A/437/375 V/118 V/120 V/121 Pereyra A.: J/ApJ/603/584 J/ApJS/141/469 Pereyra N.: J/ApJ/633/638 Perez de los Heros C.: J/other/APh/26.282 Perez E.: J/A+A/416/475 PEREZ E.: J/A+AS/93/557 Perez E.: J/AJ/110/212 Perez Hernandez F.: J/A+A/346/586 Perez I.: J/A+A/390/501 J/A+A/390/511 J/A+A/463/97 Perez M.R.: J/A+AS/104/315 J/A+AS/119/1 J/A+AS/121/223 J/A+AS/123/63 Perez-Fournon I.: J/A+A/339/34 J/A+A/467/565 J/A+A/476/1191 J/AJ/129/1198 J/AJ/129/2074 PEREZ-FOURNON I.: J/ApJS/91/507

Perez-Fournon I.: J/MNRAS/311/456 J/MNRAS/343/293 J/MNRAS/351/1290 J/MNRAS/355/97

J/MNRAS/358/333 J/MNRAS/358/397 J/MNRAS/368/1411 Perez-Gonzalez P. G.: J/A+A/365/370 Perez-Gonzalez P.G.: J/A+AS/141/409 J/ApJ/640/167 Perez-gonzalez P.G.: J/ApJ/642/126 Perez-Gonzalez P.G.: J/ApJS/154/155 J/MNRAS/338/508 J/MNRAS/338/525 Perez-Martinez R.: J/A+A/407/791 Perez-Montero E.: J/MNRAS/361/1063 Perez-Ramirez D.: J/A+A/452/739 J/MNRAS/317/234 Perez-Torres M.A.: J/MNRAS/379/275 Perie J.P.: I/230 Perie J.-P.: I/272 Perie J.P.: I/303 J/A+A/361/1143 J/A+A/438/769 J/A+AS/124/437 Perillat P.: J/AJ/130/2613 Perina S.: J/A+A/419/139

Perinotto M.: III/130 III/30A J/A+A/280/581 J/A+A/332/721 J/A+A/367/498 J/A+A/379/90

J/A+A/410/905 J/MNRAS/349/793 Perley D.: J/ApJ/645/955 Perley R. A.: J/ApJS/124/285 Perley R.: J/A+A/446/97 J/A+A/456/791 J/AJ/98/419 Perley R.A.: J/A+A/402/171 J/AJ/87/859 J/AJ/96/1655 J/ApJ/591/640 J/ApJS/74/181 VIII/65 VIII/79 Perlman E.S.: J/AJ/115/1253 J/ApJS/140/265 J/MNRAS/323/757 J/MNRAS/351/83 Perlmutter S.: J/A+A/437/789 Perola G.C.: J/A+A/409/79 J/A+A/466/31 J/ApJ/570/100 J/MNRAS/327/771 J/MNRAS/342/575 Peroux C.: J/A+A/370/426 J/A+A/410/33 J/MNRAS/346/1103 J/MNRAS/354/L25 J/MNRAS/371/495 Perova N.B.: II/139B II/140 Perrett K.M.: J/AJ/123/2490 Perri M.: J/A+A/362/799 J/A+A/434/385 J/A+A/448/L9 J/A+A/470/787 J/A+A/474/793

Perrier C.: J/A+A/331/581 J/A+A/364/665 J/A+A/410/1039 J/A+A/410/1051 J/A+A/414/351

J/A+A/474/293 J/A+A/475/1125 Perrier-Bellet C.: J/A+A/375/L27 Perrin G.: J/A+A/393/183 J/A+A/453/155 Perrin M.-N.: J/A+A/404/661 PERRY C.L.: II/100 II/138 Perry C.L.: II/88 J/PASP/108/772 Perry R.B.: J/A+A/459/875 Perryman M.A.C: I/196 Perryman M.A.C.: J/A+A/331/81 Persi P.: J/A+A/316/102 J/A+A/350/883 J/A+A/357/1020 J/A+A/357/219 J/A+A/372/173 J/A+A/376/907

J/A+A/399/995 J/A+A/401/613 J/A+A/420/945 J/A+A/421/623 J/A+A/445/971 J/MNRAS/367/513 Persic M.: J/A+A/342/671 J/ApJS/99/501 PERSIC M.: VII/181 Persson S.E.: J/AJ/109/1131 J/AJ/128/2239 J/ApJS/107/19 VII/75 Perucho M.: J/A+A/476/L17 Pesch P.: J/ApJS/66/387 PESCH P.: J/ApJS/71/549 Pesch P.: J/ApJS/86/453 Pesenson I.: J/PASP/110/779 Pesenson M.: J/AJ/131/250 Pestalozzi M.R.: J/A+A/432/737 PETERS G.: VII/18 PETERS J.: I/24 Peters J.: J/AJ/131/527 J/ApJ/510/659 J/PASP/111/438 Peters W.: J/AJ/131/527 Petersen C.: I/207 PETERSEN J.O.: J/A+AS/105/145 Peterson B.: J/ApJ/542/257 J/ApJ/598/597 J/ApJS/154/585 Peterson B. M.: J/ApJ/510/659 Peterson B.A.: J/AJ/114/326 J/AJ/127/334 J/ApJ/475/445 J/ApJ/552/289 J/ApJ/562/337 J/ApJS/107/201

J/ApJS/141/23 J/ApJS/141/31 J/MNRAS/243/1 J/MNRAS/329/227 J/MNRAS/329/87 J/MNRAS/333/100 J/MNRAS/339/652 J/MNRAS/350/1485 J/MNRAS/352/939 VIII/34

Peterson B.M.: J/A+A/369/57 J/ApJ/613/682 J/ApJ/632/799 J/ApJS/115/185 PETERSON C.J.: II/117 Peterson C.J.: J/AJ/91/530 Peterson E.R.: J/AJ/114/2822 Peterson J.R.: J/A+A/420/135 Peterson R.C.: J/AJ/124/3255 J/ApJ/420/612 J/ApJ/443/124 Peterson Ruth C.: J/ApJ/347/251 Petersson J.H.: III/234 Pethick C.J.: J/ApJ/416/276 Petit C.: J/A+AS/102/57 J/A+AS/117/467 J/A+AS/140/89 J/A+AS/144/475 J/A+AS/91/371 VII/234

VII/237 VII/238 VII/242 Petit H.: J/A+A/268/419 J/A+A/309/446 J/A+AS/131/317 J/A+AS/131/319 Petit J.-M.: J/A+A/407/1139 Petit M.: III/108 III/118 III/160A PETIT M.: III/184 Petit M.: J/A+A/344/943 J/A+AS/131/319 Petit P.: J/A+A/458/569 J/A+A/475/1053 Petitjean P.: J/A+A/283/759 J/A+A/317/416 J/A+A/376/28 J/A+A/473/791 J/A+AS/135/83 Petkie D.T.: J/ApJS/158/188 Peton-Jonas D.: J/A+AS/45/193 Petre R.: J/ApJ/613/343 J/ApJS/96/303 Petr-Gotzens M.G.: J/A+A/458/461 PETRIE R.M.: III/29 Petro L.: J/AJ/112/1335 Petrosian A.: J/A+A/434/887 PETROSIAN A.R.: J/A+A/297/49 Petrosian A.R.: J/A+AS/65/349 J/ApJS/101/287 J/MNRAS/328/1181 Petrosian G.V.: J/other/Ap/44.335 J/other/Ap/48.344 Petrosian M.B.: VII/89B VII/89B VII/89B VII/89B Petrov G.: J/A+A/376/745 Petrov L.: J/AJ/126/2562 J/AJ/129/1163 J/AJ/131/1872 J/ApJS/141/13 Petrov P.P.: J/A+A/341/553 J/A+A/409/169 PETROV P.P.: J/A+AS/107/9 Petrova A.V.: J/AJ/117/587 J/PAZh/26/301 Petry C.E.: J/AJ/128/544 J/ApJ/524/536 J/ApJS/126/133 PETTERSEN B.R.: J/A+A/300/323 Pettersen B.R.: J/MNRAS/368/1392 Petterson O.K.L.: J/MNRAS/304/733 J/MNRAS/362/1167 PETTERSSON B.: J/A+A/293/75 J/A+AS/104/233 J/A+AS/76/101 Pettersson B.: J/AJ/127/1117 Pettini M.: J/A+A/415/993 J/ApJ/519/1 J/ApJ/592/728 Petty-Powell S.: J/ApJS/140/331 Pety J.: J/A+A/475/559 Pevunova O.: J/ApJS/147/29 J/PASP/113/10 Peyaud A.E.J.: V/127 Peytremann E.: VI/10 Pezzuto S.: J/A+A/447/221 Pfefferkorn F.: J/A+A/368/797 Pfeffermann E.: IX/10A IX/29 J/A+A/316/147 Pfeifer R.H.: VI/105 Pfeiffer M.: J/ApJS/115/185 PFITSCHER K.: J/A+AS/93/553 Phan A.T.T.: J/MNRAS/366/1230 Phan-Bao N.: J/A+A/380/590 J/A+A/401/959 J/A+A/441/653 J/A+A/446/515 Phelps R.: J/ApJ/463/26 J/ApJ/507/655 Phelps R.L.: J/AJ/108/585 J/AJ/113/1723 J/ApJ/483/826 J/ApJ/508/491 J/ApJ/512/48 J/ApJ/521/155

J/ApJ/523/540 Philip A.G.D.: II/37 J/BaltA/12/323 J/BaltA/13/1 J/BaltA/15/483Short V/14 PHILIP A.G.D.: V/16 V/23 Phillipps S.: J/A+A/476/59 J/MNRAS/342/738 J/MNRAS/362/689 J/MNRAS/363/1019 V/127 Phillips A.: J/ApJ/571/136 Phillips A.C.: J/AJ/112/2408 J/ApJ/481/673 J/ApJ/599/1006 J/ApJ/625/6 J/ApJS/142/1 J/ApJS/159/41 Phillips A.J.: J/ApJS/157/175 Phillips C.: J/MNRAS/367/1609 J/MNRAS/371/898 PHILLIPS J.P.: J/A+AS/104/169 Phillips J.P.: J/A+AS/134/241 J/ApJS/139/199 J/MNRAS/353/589 J/MNRAS/357/732 Phillips K.J.H.: J/A+A/304/563 Phillips M.M.: II/179 J/AJ/112/2408 J/ApJ/594/1 Phillips P.M.: VIII/72 Phillips T.G.: J/A+A/279/541 J/ApJ/604/717 J/ApJS/156/127 PHILLIPS T.G.: VII/50 Pi C.-L. M.: J/A+AS/136/379 Pi C.-L.M.: J/A+AS/133/197 Pian E.: J/A+A/327/61 J/A+A/469/503 Piatek S.: J/AJ/121/841 J/AJ/123/2511 Piatti A.E.: J/A+A/360/529 J/A+A/370/931 J/A+A/375/30 J/A+A/379/453 J/A+A/388/179 J/A+A/399/543

J/A+A/418/979 J/A+A/421/991 J/A+A/440/111 J/A+AS/107/39 J/A+AS/127/423 J/A+AS/128/131 J/A+AS/134/301 J/AJ/116/723 J/ApJS/116/263 J/MNRAS/335/233 J/MNRAS/340/1249 J/MNRAS/346/390 J/MNRAS/349/641 J/MNRAS/358/1215 J/MNRAS/367/599 J/MNRAS/377/1737 J/MNRAS/377/300 J/other/NewA/11.262 J/PASP/117/22

Piatti: J/A+A/409/541 Pica A.J.: J/ApJS/104/287 Picat J.P.: J/A+A/421/509 J/A+A/428/1043 J/A+A/451/1159 J/A+A/457/79 J/A+A/459/679 Picaud S.: J/A+A/453/635 Piccinotti G.: J/ApJ/253/485 Pichon C.: J/A+A/376/28 PICKERING E.C.: III/1 Pickering E.C.: III/135A PICKERING E.C.: III/99 Pickering J.C.: J/A+A/377/361 J/A+A/416/655 J/A+AS/130/541 Pickles A.: J/AJ/123/2597 Pickles A.J.: J/PASJ/55/1079 VII/102 Pieloth D.: J/other/APh/26.282 Pier J.R.: I/263 J/AJ/103/638 J/AJ/105/1571 J/AJ/117/2329 J/AJ/121/2308 J/AJ/121/2331 J/AJ/121/31

J/AJ/123/567 J/AJ/126/1023 J/AJ/126/2579 J/AJ/127/2838 J/AJ/127/3034 J/AJ/87/1515 J/ApJ/607/426 J/ApJS/53/791

Pierce M.J.: J/AJ/112/2471 J/AJ/126/2644 J/ApJ/533/744 J/ApJS/80/479 J/MNRAS/362/689 V/127 Pierce-Price D.: J/MNRAS/368/1223

Pierini D.: J/A+A/361/863 J/A+A/414/905 J/A+AS/120/489 J/A+AS/120/521 J/A+AS/121/507

J/A+AS/125/293 J/A+AS/137/101 J/A+AS/142/65 J/A+AS/142/73 J/MNRAS/332/422 Pierre M.: J/A+A/363/440 J/A+A/392/795 J/A+A/439/413 J/A+A/456/791 J/A+A/467/73 J/A+A/471/1105

J/A+A/474/473 J/A+AS/124/283 J/ApJ/591/640 J/MNRAS/382/279 Piersimoni A.: J/A+AS/125/399 PIERSIMONI A.M.: J/A+AS/101/195 Piersimoni A.M.: J/AJ/121/916 J/AJ/129/2714 Piestrzynska B.: VIII/82 Piestrzynski B.: VIII/70A VIII/78 VIII/81 Pietila H.: J/A+AS/132/305 J/A+AS/143/357 Pietrinferni A.: J/A+A/432/851 Pietrukowicz P.: J/MNRAS/365/110 Pietrzynski G.: II/278 J/A+A/371/497 J/A+A/440/783 J/A+A/450/873 J/A+A/478/659 J/AcA/48/147

J/AcA/48/175 J/AcA/48/563 J/AcA/49/223 J/AcA/49/319 J/AcA/49/437 J/AcA/49/521 J/AcA/49/543 J/AcA/50/1 J/AcA/50/337 J/AcA/50/355 J/AcA/51/175 J/AcA/51/221 J/AcA/51/317 J/AcA/52/129 J/AcA/52/143 J/AcA/52/217 J/AcA/53/1 J/AcA/53/93 J/AcA/54/347 J/AJ/123/789 J/AJ/128/1167 J/AJ/128/2815 J/ApJ/636/240 J/MNRAS/348/1439 J/MNRAS/350/679 J/MNRAS/356/331 J/MNRAS/361/1055 J/PASP/114/298

Pietsch W.: II/208 IX/10A IX/11 IX/29 J/A+A/278/36 J/A+A/305/74 J/A+A/317/328 J/A+A/318/111 J/A+A/344/521 J/A+A/361/491 J/A+A/373/438 J/A+A/373/63 J/A+A/378/30 J/A+A/426/11 J/A+A/434/483 J/A+A/448/1247 J/A+A/451/835 J/A+A/465/375 J/A+AS/114/215 J/A+AS/122/201 J/A+AS/127/119 J/A+AS/136/81 J/A+AS/139/277 J/A+AS/142/41 J/A+AS/143/391 J/A+AS/147/75

Pietsh W.: J/A+AS/132/341 Pignata G.: J/AJ/131/1648 J/MNRAS/369/1880 Pignatelli E.: J/A+A/470/39 Pigulski A.: J/A+A/325/987 J/A+A/345/505 J/A+A/376/144 J/A+A/398/541 J/A+A/434/89 J/A+A/477/907

J/A+A/477/917 J/AcA/53/27 J/MNRAS/298/753 Piirola V.: J/A+A/454/277 PIIROLA V.: J/A+AS/102/343 Piironen J.: J/A+A/374/1049 Pike C.D.: J/MNRAS/274/1194 PILACHOWSKI C.A.: J/AJ/105/2299 Pilachowski C.A.: J/AJ/112/545 J/AJ/113/279 J/AJ/129/2725 J/PASP/111/1233 Pilcher F.: J/A+A/465/331 Pildis R.A.: J/ApJS/126/209 Pilecki B.: II/264 Pilippenko A. V.: J/ApJ/510/659 Pilkington J.D.H.: J/A+A/312/1031 PILKINGTON J.D.H.: VII/53 VIII/4 Pillai T.: J/A+A/434/613

Pillitteri I.: J/A+A/399/919 J/A+A/412/495 J/A+A/421/175 J/A+A/426/1007 J/A+A/430/287

J/A+A/450/993 J/A+A/456/977 J/A+A/457/501 J/A+A/462/123 J/A+A/468/353 Pimbblet K.: J/MNRAS/372/425 J/MNRAS/375/68 Pimbblet K.A.: J/MNRAS/327/588 J/MNRAS/366/645 J/MNRAS/368/651 Pimenova T.: J/ApJ/626/L5 Pindao M.: J/A+A/355/89 J/A+AS/136/35 Pindor B.: J/AJ/131/1 J/AJ/131/49 Pineau des Forets G.: J/A+A/321/293 J/A+A/356/1010 Pineault S.: J/MNRAS/246/169 Pineda P.: J/ApJ/621/104 Piner B.G.: J/ApJS/155/33 J/PASJ/52/997 Pinfield D.J.: J/A+A/421/763 J/MNRAS/313/347 J/MNRAS/342/1241 J/MNRAS/372/163 Pinheiro F.J.G.: J/A+A/399/271 Pinkau K.: J/ApJS/94/551 Pinkney J.: J/AJ/120/2269 J/AJ/129/2138 J/ApJ/416/36 Pinsonneault M.H.: J/ApJ/602/L117

Pintado O.I.: J/A+A/354/899 J/A+A/367/859 J/A+A/371/L5 J/A+A/406/987 J/A+A/412/721 J/A+A/443/157 J/A+A/454/171 J/A+A/459/871 J/A+A/462/591 J/A+AS/125/219 J/A+AS/129/563 J/MNRAS/362/1025

Pinte C.: J/A+A/463/1017 J/AJ/127/2723 Piotto G: J/A+A/315/40 Piotto G.: J/A+A/333/479 J/A+A/391/945 J/A+A/417/597 J/A+A/432/851 J/A+A/434/235 J/A+A/436/527

J/A+A/440/901 J/A+A/444/831 J/A+A/459/489 J/A+A/464/939 J/A+A/464/953 J/A+A/471/515 J/A+AS/144/5 J/A+AS/145/451

Piquard S.: II/233 J/A+A/373/576 Piranomonte S.: J/A+A/434/385 J/A+A/470/787 Piro L.: J/A+A/455/813 Pirogov L.: J/A+AS/133/337 PIRONI R.: II/155 Pisani A.: J/A+A/342/1 J/A+A/470/39 J/AJ/113/483 J/AJ/123/2976 J/ApJ/543/178

Piskunov A.E.: III/239 III/254 J/A+A/320/79 J/A+A/332/575 J/A+A/384/145 J/A+A/438/1163 J/A+A/440/403 J/A+A/468/151 J/A+A/477/165 J/A+AS/134/525 J/AN/325/740 J/AN/328/889 J/AZh/80/291 V/19

Piskunov N.: J/A+A/374/265 J/A+AS/142/467 Piters A.J.M.: J/A+A/325/1115 J/A+AS/118/545 J/A+AS/128/29 Pitman K.M.: J/MNRAS/371/1744 Pittard J.M.: J/A+A/395/499 Pittella G.: J/A+AS/94/327 PITTS P.S.: II/184 Pitts P.S.: II/209 II/216 II/225 Pivalica S.: J/A+AS/102/607 Pizagno J.: J/ApJ/633/844

Pizzella A.: J/A+A/342/671 J/A+A/360/439 J/A+A/374/394 J/A+A/382/488 J/A+A/408/873

J/A+A/416/507 J/A+A/424/447 J/A+A/440/107 J/AJ/119/2540 Pizzella: J/A+AS/120/257 Pizzichini G.: J/other/GCN/1070.1 Pizzolato N.: J/A+A/397/147 J/A+A/410/671 Plachinda S.: J/A+A/469/1069 Plaga R.: J/A+A/390/39 PLAIN Center (2001): VI/117 Plambeck R.L.: J/ApJS/149/343 Plana H.: J/AJ/130/1707 Planesas P.: J/ApJS/160/272 PLATAIS I.: I/177 I/178 Platais I.: I/203 PLATAIS I.: I/212 Platais I.: I/277 I/288 PLATAIS I.: III/140 Platais I.: J/A+A/397/997 J/A+A/399/1167 J/A+A/403/993 J/A+A/461/509 J/AJ/113/1045 J/AJ/119/2866

J/AJ/122/1486 J/AJ/122/1586 J/AJ/124/2254 J/AJ/124/601 J/AJ/126/2922 Platania P.: J/A+A/397/213 J/A+A/410/847 Platt D.: J/ApJS/83/203 Pleticha D.: VIII/61 Plez B.: J/A+A/333/231 J/A+A/379/305 J/A+A/403/1105 J/A+A/416/1117 J/A+A/457/637 PLEZ B.: J/A+AS/105/311 Plionis M.: J/ApJ/639/37 J/ApJS/95/401 J/MNRAS/349/135 Plugina A.I.: I/44 Pluzhnik E.A.: J/A+A/385/87 J/A+A/409/245 J/A+A/422/627 J/A+AS/105/503 J/A+AS/140/287 Plyasheshnikov A.: J/A+A/390/39 J/A+A/406/L9 J/A+A/410/813 Pocock S.B.: V/32A V/4 Podesta R.C.: J/A+A/430/327 J/A+AS/118/1 J/A+AS/136/1 J/other/PBeiO/32.87 Podsiadlowski P.: J/MNRAS/380/1098 Poeppel W.G.L.: VIII/75 VIII/76 VIII/9 Poetzi W.: J/A+A/447/735 Pogge R. W.: J/ApJ/510/659 Pogge R.W.: J/A+A/369/57 J/AJ/124/3042 J/ApJ/517/650 J/ApJ/613/682 J/ApJ/632/799 J/ApJS/143/73

J/ApJS/146/353 J/ApJS/162/80 Poggianti B.: J/A+A/444/365 J/ApJ/630/206

Poggianti B.M.: J/A+A/382/60 J/A+A/427/397 J/A+A/470/39 J/A+AS/122/399 J/ApJ/542/673

J/ApJS/122/51 J/ApJS/137/279 J/MNRAS/371/1777 Pohjolainen S.: J/A+AS/113/115 Pohl A.C.: J/other/APh/26.282 Pohl B.L.: J/AJ/122/1586 Pohl M.: J/A+AS/131/11 J/ApJS/107/227 J/ApJS/123/79 Pohlen M.: J/A+AS/145/405 Pohlman J.W.: I/268 Poirier S.: J/A+A/427/397 J/A+A/444/365

Pojmanski G.: II/264 II/264 II/264 II/264 II/264 J/A+A/477/907 J/A+A/477/917 J/AcA/48/35

J/AcA/50/177 Pokorny R.S: J/A+A/397/575 Pokorny R.S.: J/A+A/421/763 Polak A.A.: J/AJ/128/245 Polatidis A.G.: J/ApJS/98/1 J/ApJS/98/33 J/ApJS/99/297 Polcaro V.F.: J/A+A/458/225 Polehampton E.T.: J/A+A/406/L47 Poli F.: J/A+A/375/1 J/A+A/392/395 J/AJ/120/2206 Poljakov E.V.: J/A+A/396/125 Polk K.: J/ApJS/112/557 Pollacco D.: J/A+A/467/785 J/MNRAS/362/1006 Pollacco D.L.: J/A+A/414/23 J/A+A/452/503 J/MNRAS/362/117 J/MNRAS/372/1117 J/MNRAS/379/647

J/MNRAS/379/816 Pollack L.K.: J/ApJ/589/733 POLLACO D.L. Astron. Astrophys. Suppl. Ser. 94: J/A+AS/94/73 Pollanen M.D.: J/PASP/107/617 Pollard D.L.: J/ApJ/628/729 Pollard K.R.: J/MNRAS/331/609 J/MNRAS/349/833 Polletta M.: J/A+A/467/73 J/A+A/474/473 J/AJ/129/1198 J/ApJ/633/871 J/ApJS/106/399 Polletta M.C.: J/A+A/467/565 Polletta M.D.C.: J/AJ/129/2074 Pollmann E.: J/A+A/408/305 Pollo A.: J/A+A/403/857 J/A+A/428/1043 J/A+A/457/79 Pollock A.M.T.: J/A+AS/134/287 J/MNRAS/375/145 Pollock J.T.: J/A+A/356/L21 Polojentsev D.D.: I/195 I/242 POLOJENTSEV D.D.: I/43 Polojentsev D.D.: I/82 POLOJENTSEVA T.A.: I/43 Polojentseva T.A.: I/82 Polomski E.: J/ApJS/95/1 Polosukchina N.S.: J/A+AS/105/125 POLOZHENTSEV D.D.: I/159A Polozhentsev D.D.: I/42 I/44 Polozhentseva T.A.: I/44 Polyakov E.V.: J/PAZh/27/456 Pomerleau C.: J/ApJS/144/47 Pompei E.: J/A+A/423/849 J/A+A/445/857 J/A+A/473/399 Pompilio F.: J/ApJ/570/100 Ponce V.C.: J/AJ/129/2798 Poncelet J.M.: V/53A Poncy R.: J/A+A/465/331 Ponman T.J.: J/MNRAS/274/1194 J/MNRAS/328/461 J/MNRAS/382/279 Ponnampalam T.: J/AJ/122/2587 Ponomareva G.A.: II/140 Pont F.: J/A+A/318/416 J/A+A/341/818 J/A+A/371/497 J/A+A/424/L31 J/A+A/466/743 J/A+A/471/L51

J/A+A/476/1347 J/AJ/123/789 J/AJ/127/840 V/117 Pooley D.M.: J/MNRAS/298/637 Pooley G.G.: J/A+A/407/1039 J/A+A/447/245 J/MNRAS/308/473 J/MNRAS/341/1066 J/MNRAS/342/915

J/MNRAS/354/485 J/MNRAS/360/340 J/MNRAS/379/1442 VIII/32 Pope A.: J/AJ/121/2308 J/AJ/121/2331 J/AJ/123/567 J/MNRAS/358/149 J/MNRAS/372/1621 Popescu C.C.: J/A+AS/116/43 J/A+AS/135/511 J/A+AS/142/399 J/ApJ/619/755 Popescu M.D.: J/A+AS/144/481 Popesso P.: J/A+A/423/449 J/A+A/454/423 J/A+A/461/397 J/A+A/478/83 Popov M.: J/PASJ/52/997 Popova A.A.: J/A+A/323/113 POPOVA E.: V/41 Popovic G.M.: J/A+AS/123/487

Popovic L.C.: J/A+A/365/656 J/A+A/399/795 J/A+A/417/375 J/A+AS/116/359 J/A+AS/120/373

J/A+AS/127/295 J/A+AS/128/203 Popowski P.: J/AJ/127/334 J/ApJ/542/257 J/ApJ/552/289 J/ApJ/562/337 J/ApJ/598/597 Poppel: J/A+AS/142/25 POPPER D.M.: V/63 Poretti E.: J/A+A/319/867 J/A+A/371/986 J/A+A/374/1056 J/A+A/450/715 J/A+A/471/515 PORETTI E.: J/A+AS/108/55 Poretti E.: J/A+AS/133/7 Porquet D.: J/A+A/376/1113 Porras A.: J/A+A/361/660 Porras A: J/A+A/399/1135 Porras A.: J/AJ/126/1916 Porrata R.: J/other/APh/26.282 Portegies Zwart S.F.: J/A+A/430/137 J/AJ/124/1570 PORTER A.C.: J/AJ/101/1561 Porter A.C.: J/ApJ/462/614 J/ApJS/104/185 Portinari L.: J/A+A/335/823 J/MNRAS/373/13 Porto de Mello G.F.: J/A+A/434/275 J/A+A/441/1149 J/A+A/458/113 Posselt B.: J/ApJS/161/361 Possenti A.: J/MNRAS/328/17 J/MNRAS/335/275 J/MNRAS/342/1299 J/MNRAS/352/1439 J/MNRAS/372/777 Posson-Brown J.: J/AJ/129/1311

Postman M.: J/AJ/112/1335 J/AJ/116/560 J/AJ/116/584 J/AJ/120/1 J/AJ/122/1125 J/AJ/125/478

J/AJ/127/131 J/AJ/128/1990 J/AJ/130/968 J/AJ/131/143 POSTMAN M.: J/ApJ/365/66 J/ApJ/384/404 Postman M.: J/ApJ/425/418 J/ApJ/440/28 J/ApJ/569/689 J/ApJ/579/93 J/ApJ/612/202 J/ApJ/621/53

J/ApJ/644/30 J/ApJ/661/750 J/ApJS/148/243 J/ApJS/150/1 Postnov K.A.: J/PAZh/30/589 Posztobanyi K.: J/A+A/430/1049 J/A+A/445/617 POTEKHIN A.Y.: J/A+AS/104/89 Pott J.-U.: J/A+A/433/117 Pottach S.R.: J/A+AS/126/297 Pottasch S.R.: J/A+A/274/895 POTTASCH S.R.: J/A+A/299/238 Pottasch S.R.: J/A+A/458/173 J/A+AS/126/479 Potter H.I.: I/195 POTTER K.H.: I/159A Pottschmidt K.: J/A+A/407/1039 J/A+A/447/245 POULAIN P.: J/A+AS/103/573 Poulain P.: J/A+AS/116/429 J/A+AS/126/67 J/A+AS/141/113 Poulet F.: J/A+A/407/1139 Poulsen M.A.: J/A+A/371/79 Pounds K.A.: J/MNRAS/260/77 VII/19 Pourbaix D.: B/sb9 J/A+A/397/997 J/A+A/399/1167 J/A+A/403/993 J/A+A/423/755 J/A+A/442/365

J/A+A/444/643 J/AJ/129/1001 V/122 Poutanen J.: J/ApJ/563/80 Poveda: J/A+A/356/529 Power J.: J/A+A/475/1053

Pozzetti L.: J/A+A/371/833 J/A+A/392/395 J/A+A/403/857 J/A+A/428/1043 J/A+A/432/69

J/A+A/437/883 J/A+A/457/79 J/A+A/465/1099 Pozzi A.: II/224

Pozzi F.: J/A+A/369/380 J/AJ/127/3075 J/MNRAS/325/1173 J/MNRAS/351/1290 J/MNRAS/358/397

J/MNRAS/371/1891 Pozzo M.: J/MNRAS/335/291 J/MNRAS/341/805 Prabhu T.P.: J/other/BASI/30.911 PRADA F.: J/A+A/293/703 Prada F.: J/AJ/127/1811 J/AJ/127/704 J/ApJ/598/260 J/ApJ/633/844 J/ApJ/645/228 J/ApJS/141/61 Prada: J/MNRAS/326/959

Pradhan A.K.: J/A+A/283/319 J/A+A/293/953 J/A+A/408/789 J/A+A/437/345 J/A+AS/119/509

J/A+AS/123/575 J/A+AS/126/373 J/A+AS/135/347 J/A+AS/136/395 J/A+AS/144/141 J/A+AS/147/111 PRADHAN A.K.: VI/80 PRADO P.: J/A+AS/106/339 J/A+AS/111/85 Prado P.: J/A+AS/116/75 J/A+AS/124/13 J/A+AS/139/69 PRAGER R.: I/46 Prahl J.: J/A+A/390/39 Pramskij A.: J/A+A/419/469 Pramskij A.G.: J/A+A/442/109 J/AJ/127/704 J/ApJS/153/429 Pramsky A.G.: J/A+A/366/771 J/A+A/374/907 J/A+A/397/463 Prandoni I.: J/A+A/385/39 J/A+A/415/549 J/A+A/445/465 J/A+A/457/517 J/A+A/462/553 J/A+AS/137/75

J/AJ/130/1373 VIII/63 Prato L.: J/ApJ/636/932 Pratt M.: J/ApJ/542/257 J/ApJ/598/597 Pratt M.R.: J/AJ/114/326 J/ApJ/552/289 J/ApJ/562/337 Pravdo S.: J/AJ/127/2659 J/AJ/128/787 Pravdo S.H.: J/AJ/120/1426 J/ApJ/574/258 J/ApJ/605/259 J/ApJ/626/272 Pray D.: J/A+A/465/331 Predehl P.: IX/10A IX/29 Preece R.D.: IX/20A J/ApJS/126/19

Preibisch T.: J/A+A/333/619 J/A+A/401/543 J/A+A/409/147 J/A+AS/137/305 J/AJ/121/1040

J/AJ/122/866 J/AJ/124/404 J/ApJS/160/319 Preite-Martinez A.: J/A+A/368/267 Prescott M.K.M.: J/ApJ/644/100 Pressberger R.: J/A+AS/128/573 Preston G.W.: J/AJ/103/1987 J/AJ/103/267 J/AJ/108/538 J/AJ/132/85 J/AJ/90/2089 J/ApJS/76/1001 Preston R.A.: J/PASJ/52/997 Pretz J.: J/other/APh/26.282 Preuss O.: J/A+A/409/275 J/A+A/411/565 Prevot L.: I/196 PREVOT L.: III/105 III/111 III/147 Prevot L.: J/A+A/303/137 J/A+A/319/881 J/A+A/389/149 PREVOT L.: J/A+AS/109/447 Prialnik D.: J/MNRAS/313/641 Pribulla T.: V/119 Price A.: J/other/IBVS/5526 Price P.B.: J/other/APh/26.282 Price R.M.: J/AJ/123/873 J/AJ/124/1954 J/AJ/124/690 J/AJ/128/16 J/MNRAS/337/641 J/MNRAS/346/787

J/MNRAS/355/1303 Price S.: J/ApJ/552/289

Price S.D.: II/269 II/94 III/231 J/A+A/453/535 J/AJ/122/1844 J/AJ/124/2990 J/AJ/126/1423 J/AJ/131/612 J/AJ/132/1445 J/ApJS/119/141 J/ApJS/150/343 J/ApJS/151/299 J/ApJS/154/673 V/107 V/114 V/98

Price T.: J/AJ/121/1872 Priddey R.S.: J/MNRAS/372/1621 PRIEBE A.: VII/161 Prieto C.: J/A+AS/143/335 J/AJ/127/1181 Prieto C.A.: J/MNRAS/340/304 J/MNRAS/367/1329 Prieto C.M.: J/AJ/111/936 J/AJ/119/3084 Prieto G.: J/A+A/474/121 Prieto J.L.: J/AJ/130/2453 J/AJ/131/1648 PRIETO M.: J/A+AS/93/557 Prieto M.: J/ApJS/154/155 Prieur J.L.: J/A+A/367/865 Prieur J.-L.: J/A+AS/134/545 J/ApJS/131/561 J/MNRAS/367/1170 Prigozhin G.: J/other/GCN/1070.1 Primack J.: J/AJ/128/163 Primas B.: J/A+A/403/1105 PRIMAS F.: J/A+A/290/885 Primas F.: J/A+A/390/235 J/A+A/416/1117 J/A+A/442/961 J/A+A/459/423 Primini F.A.: IX/22 J/ApJ/577/738 J/ApJ/609/735 J/ApJ/610/247 Primnini F.A.: IX/13 Prinja R.K.: J/MNRAS/284/265 Prins S.: J/A+AS/114/215 J/A+AS/126/401

Prisinzano L.: J/A+A/404/927 J/A+A/417/945 J/A+A/430/941 J/A+A/462/123 J/A+A/462/245

J/A+A/475/539 PRITCHARD J. D.: J/A+AS/109/447 Pritchard J.D.: J/A+A/303/137 J/A+A/392/215 J/MNRAS/331/609 J/MNRAS/349/833 Pritchet C.J.: J/AJ/121/1840 J/AJ/130/2043 J/ApJ/565/208 VII/208 Pritzl B.: J/AJ/123/3216 Pritzl B.J.: J/AJ/124/1464 J/AJ/127/318 J/AJ/129/2232 J/AJ/130/2140 Probst R.: J/AJ/107/2120 Prochaska J.X.: J/A+A/445/93 J/AJ/123/2206 J/ApJ/617/718 J/ApJ/635/123 J/ApJ/639/766 J/ApJ/643/680

J/PASP/116/622 J/PASP/118/1077 Prochaska L.C.: J/AJ/130/2666 Prochter G.E.: J/ApJ/639/766 J/PASP/118/1077 Proctor D.D.: J/ApJS/135/227 Produit N.: J/A+A/467/585 J/ApJS/170/175 Proffitt C.R.: J/ApJ/612/481 PROFFITT C.R.: VI/55 Promislov V.G.: J/AZh/79/610 Pronik V. I.: J/ApJ/510/659 Pronik V.I.: J/A+A/320/405 J/A+A/324/904 J/A+A/369/57 J/A+A/371/79 J/ApJS/115/185 J/other/Ap/48.304 Pronik V.I: J/PAZh/23/409 Prosser C.: J/AJ/112/2408 Prosser C.F. Bresolin F.: J/ApJ/514/614 Prosser C.F.: J/A+A/341/751 PROSSER C.F.: J/AJ/105/1441 J/AJ/106/1080 Prosser C.F.: J/AJ/110/1229 J/AJ/111/1193 J/AN/319/201 J/AN/319/215 J/ApJ/507/655 J/ApJ/512/48

J/ApJ/521/155 J/ApJS/91/625 V/80 Protitch-Benishek V.: J/A+AS/125/399 Proust D.: J/A+A/258/243 J/A+A/342/1 J/A+A/377/428 J/A+A/407/31 J/A+A/445/819 J/A+A/465/1099

J/A+AS/100/583 J/A+AS/114/565 J/A+AS/125/247 J/A+AS/130/323 PROUST D.: J/MNRAS/259/233 Proust D.: J/other/PASA/16.113 J/other/PASA/18.232 Prouton O.: J/MNRAS/358/397 Provost J.: J/A+A/478/197 Prsa A.: J/A+A/463/1061 J/A+A/464/263 Prugniel: J/A+A/384/371 J/A+AS/136/509 J/A+AS/139/483 Prugniel P.: III/218 III/251 PRUGNIEL P.: J/A+A/297/28

Prugniel P.: J/A+A/309/749 J/A+A/366/68 J/A+A/423/833 J/A+A/464/L21 J/A+AS/122/521 J/A+AS/126/15 J/A+AS/126/519 J/A+AS/127/117 J/A+AS/131/287 J/A+AS/132/255 J/A+AS/136/519 J/A+AS/140/89 J/A+AS/144/475 J/A+AS/145/263 VII/206 VII/237

Prusti T.: J/A+A/272/235 J/A+A/350/883 J/A+A/357/219 J/A+A/372/173 J/A+A/420/945 J/A+A/421/623

Pryor C.: J/AJ/107/2067 J/AJ/110/1699 J/AJ/110/2131 J/AJ/113/1026 J/AJ/115/592 J/AJ/119/1268

J/AJ/121/841 J/AJ/123/2511 J/ApJS/120/265 Przewozny D.: J/A+AS/125/399 Przybilla N.: J/A+A/437/467 Przybylski G.T.: J/other/APh/26.282 P.SIVAGNANAM: J/A+A/297/494 Ptak A.: J/ApJ/582/615 Ptak A.F.: J/ApJ/602/231 J/ApJS/143/25 PTOLEMY C.: V/61 Puccetti S.: J/A+A/457/501 J/A+A/466/31 J/A+A/469/1211 Puche D.: J/AJ/116/2341 Puchnarewicz E.M.: J/MNRAS/311/456 Puddu E.: J/A+A/428/339 Puech M.: J/A+A/455/107 J/A+A/465/1099 Puehlhofer G.: J/A+A/390/39 J/A+A/406/L9 J/A+A/410/813 PUEL F.: I/74 Puerari I.: J/A+A/379/54 J/A+A/438/507 J/AJ/124/3157 J/AJ/129/682 Puetter R.: J/A+AS/147/5 Puetter R.C.: J/A+A/459/875 J/AJ/121/362 J/ApJ/538/493 Puga E.: J/ApJS/155/123 Puget J.L.: J/A+A/350/883 J/A+A/357/219 J/A+A/372/173 Puget J.-L.: J/A+A/372/364 Puget J.L.: J/A+A/420/945 J/A+A/421/623 Puget J.-L.: J/MNRAS/302/222 J/MNRAS/305/297 J/MNRAS/351/1290 VIII/66 Puliaev S.: J/A+AS/135/227 J/A+AS/136/531 J/A+AS/143/265 Puliaev S.P.: J/A+A/384/650 J/A+AS/117/335 Pulone L.: J/A+A/330/453 J/A+A/380/478 J/AJ/109/663 J/MNRAS/299/834 Puls J.: J/A+A/321/531 J/A+A/437/467 J/ApJS/161/154 Pun C.S.J.: J/ApJS/99/223 Pundure I.: III/227 Punsly B.: J/A+A/370/468 Purcell C.R.: J/A+A/474/891 J/MNRAS/366/1096 J/MNRAS/367/1609 Purgathofer A.: J/A+A/452/959 Puric J.: J/A+AS/102/607 Pursimo T.: J/A+AS/130/109 J/A+AS/132/305 J/A+AS/143/357 Purvis A.: J/MNRAS/229/589 Puschell J.J.: J/ApJ/252/102 Puss A.: J/A+A/415/273 Pustilnik S.: J/A+A/419/469

Pustilnik S.A.: J/A+A/366/771 J/A+A/373/24 J/A+A/374/907 J/A+A/442/109 J/A+AS/131/295

J/A+AS/135/511 J/A+AS/137/299 J/A+AS/139/1 J/A+AS/142/417 J/AJ/127/704 J/ApJS/153/429 Pustilnik S.A. Kniazev A.Y.: J/A+A/397/463 Putman M.: J/MNRAS/355/1303 Putman M.E.: J/AJ/123/873 J/AJ/124/1954 J/AJ/124/690 J/AJ/128/16 J/ApJS/142/71 J/ApJS/165/307

J/MNRAS/337/641 Puzakova T.Yu.: J/PAZh/32/628 J/PAZh/32/672 Puzeras E.: J/A+A/431/933 Puzia T.H.: J/A+A/361/491 J/A+A/415/123 J/A+A/442/85 J/MNRAS/372/1259 Pych W.: J/A+A/424/1101 J/A+A/465/331 J/A+AS/126/479 J/AJ/124/1486 J/AJ/130/767 Pye J.P.: J/A+A/476/1191 J/MNRAS/274/1165 J/MNRAS/326/1161 J/MNRAS/367/781 VII/19 Pyke R.: J/AJ/110/2715 Pynzar A.V.: J/AZh/71/110 Pyper D.M.: III/89 J/A+A/337/183 Python M.: II/14

Qian B.-C.: J/A+A/369/57 Qian B.C.: J/A+A/369/758 Qian B.-C.: J/ApJS/115/185 Qian S.: J/A+A/355/171 J/A+AS/125/475 Qian S.-B.: J/AJ/130/2252 J/AJ/131/1032 Qian Y.: J/ApJ/590/73 Qiao Q.Y.: J/A+A/430/327 Qiao R.C.: J/A+A/391/775 J/A+A/422/377 J/A+A/454/379 J/MNRAS/376/1707 Qin S.: J/A+A/426/503 Qiu H.M.: J/A+A/397/257 J/A+A/425/671 J/ApJS/126/461 Qiu Y.: J/MNRAS/369/1880 Qiu Y.-H.: J/other/AcApS/16.103 Qiu Y.L.: J/AJ/130/1640 Qu Y.: J/ApJ/562/528 Quadri R.: J/AJ/134/1103 J/ApJS/162/1 Quaintrell H.: J/A+A/441/685 Qualtrough C.J.: J/A+A/440/151 Quast G.: J/A+AS/133/299 Quast G.R.: J/A+A/460/695 J/A+AS/128/561 J/AJ/126/2971 Queinnec F.: J/A+A/303/137 QUEINNEC F.: J/A+AS/109/447 Queloz D.: J/A+A/318/416 J/A+A/335/183 J/A+A/356/590 J/A+A/375/205 J/A+A/375/L27 J/A+A/379/999

J/A+A/388/632 J/A+A/390/267 J/A+A/392/215 J/A+A/406/373 J/A+A/410/1039 J/A+A/410/1051 J/A+A/414/351 J/A+A/415/391 J/A+A/458/997 J/A+A/466/743 J/A+A/471/L51 J/A+A/474/293 J/A+A/475/1125 J/A+A/476/1347

QUERCI F. (France): VI/26 Quick J.: J/PASJ/52/997 Quick J.F.H.: J/AJ/127/3609 J/AJ/130/2529 Quillen A.C.: J/ApJ/632/799 J/ApJS/143/73 J/ApJS/149/327 Quilligan F.: J/A+A/385/377 Quimby R.: J/A+A/437/789 QUINET P.: J/A+AS/102/435 Quinet P.: J/A+AS/116/573 J/A+AS/119/99 J/A+AS/120/361 J/A+AS/129/147 J/A+AS/129/603 Quiniento Z.M.: J/A+AS/97/435 Quinn D.P.: J/MNRAS/369/97 Quinn P.: J/ApJ/542/257 J/ApJ/598/597 Quinn P.J.: J/AJ/114/326 J/AJ/127/334 J/ApJ/552/289 J/ApJ/562/337 Quinn T.: J/MNRAS/351/265 Quinney S.J.: J/AJ/128/1558

Quintana H.: J/A+A/258/243 J/A+A/377/428 J/A+A/433/757 J/A+A/445/819 J/A+AS/101/475 J/A+AS/119/329 J/A+AS/120/357 J/A+AS/125/247 J/A+AS/134/59 J/AJ/110/463 J/AJ/112/36 J/AJ/129/1237 J/AJ/130/2012 J/ApJ/502/558 J/ApJ/562/254 J/ApJ/594/154 J/ApJS/126/1 J/ApJS/153/397 J/ApJS/96/343 J/MNRAS/382/279 J/other/PASA/16.113

Quireza C.: J/A+A/345/629 J/A+A/475/217 J/ApJS/165/338

Quirrenbach A.: J/A+A/365/110 J/A+A/378/116 J/A+A/379/564 J/A+A/458/461 J/AJ/123/2206 J/ApJ/439/983 J/ApJ/466/254 J/MNRAS/358/795

R.: J/A+A/374/504 J/A+A/385/1014 J/A+A/395/753 J/A+A/475/915 J/A+AS/139/483 J/A+AS/142/25 Raassen A.J.J.: J/A+A/340/300 J/A+A/376/1113 J/A+AS/125/539 J/A+AS/130/541 Rabbia Y.: J/A+AS/110/107 J/A+AS/134/545 Rabinowitz D.: J/AJ/127/1158 Raboud D.: J/A+A/315/384 J/A+A/325/167 J/A+A/329/101 J/A+A/333/897 J/A+A/360/509 Rachkovskaya T.M.: J/AZh/79/841 Racine M.: J/AJ/109/1055 Racioppi F.: J/A+AS/121/119 Racusin J.: J/A+A/448/L9 Radbo M.: J/other/IBVS/5633 Radburn-Smith D.J.: J/MNRAS/369/1131 Radecke H.-D.: J/A+A/319/18 J/ApJS/101/259 J/ApJS/94/551 Radocinski R.G.: J/ApJS/127/79 Radovich M.: J/A+A/403/857 J/A+A/428/1043 J/A+A/428/339 J/A+A/439/413 J/A+A/457/79 Rafal’skii V.B.: J/PAZh/27/456 Rafanelli P.: J/A+A/368/797 Rafelski M.: J/AJ/129/2701

Rafferty T.J.: I/268 J/AJ/119/3071 J/AJ/120/1120 J/AJ/121/3224 J/AJ/122/1586 J/AJ/124/2254

J/AJ/125/2728 J/AJ/127/539 J/AJ/128/3012 Raftery A.: J/ApJ/508/314 Raftery C.L.: J/MNRAS/374/399 Raga A.C.: J/AJ/130/1707 Ragaini S.: J/A+A/434/1107 Ragone C.J.: J/A+A/445/819 Raguzova N.V.: J/AZh/80/535 Rahal Y.R.: J/A+A/454/185 Raharto M.: J/A+AS/114/487 J/A+AS/134/87 Raimondo G.: J/A+A/438/521 J/A+A/462/107 Raines S.N.: J/ApJ/639/816

Raiteri C.M.: J/A+A/327/61 J/A+A/395/11 J/A+A/402/151 J/A+AS/121/119 J/A+AS/127/445

J/A+AS/129/577 J/A+AS/130/109 J/A+AS/132/361 J/A+AS/144/481 Rakhaminov Sh.Yu.: IX/8 Rakos K.: J/AJ/124/3061 J/AJ/130/1002 Rakos K.D.: J/A+A/403/937 Ralite N.: I/272 III/200B III/221A RAM SAGAR JOSHI U.C.: V/43 Ramanamurthy P.V.: J/ApJS/101/259 J/ApJS/107/227 Ramaprakash A.N.: J/A+AS/141/175 Ramdani A.: J/A+A/372/85 Ramella M.: J/A+A/312/745 J/A+A/342/1 J/A+A/416/839 J/A+A/442/841 J/A+A/444/685 J/A+A/470/39

J/A+AS/130/323 J/AJ/105/393 J/AJ/109/1458 J/AJ/113/483 J/AJ/116/1 J/AJ/123/2976 J/ApJ/518/69 J/ApJ/534/114

Ramirez A.: J/AJ/110/463 J/AJ/112/36 J/ApJ/562/254 J/ApJS/96/343 Ramirez I.: J/ApJ/626/446 Ramirez: J/A+A/465/271 Ramirez S.: J/AJ/132/137 Ramirez S.V.: J/AJ/127/2659 J/AJ/128/787 J/ApJ/597/323 J/ApJ/607/904 J/ApJS/143/73 Ramirez-Ruiz E.: J/ApJ/634/L89 J/ApJ/641/L13 Rampazzo: J/A+AS/139/483 RAMPAZZO R.: J/A+A/297/28

Rampazzo R.: J/A+A/330/423 J/A+A/345/419 J/A+A/433/497 J/A+A/445/79 J/A+A/463/455

J/A+AS/110/131 RAMPAZZO R.: J/A+AS/110/279 Rampazzo R.: J/A+AS/130/251 J/AJ/129/1832 J/MNRAS/364/146 J/other/ApL/30.1 RAMSAUER J.: J/A+AS/113/71 Ramsay G.: J/A+A/459/777 RAMSAY G.: J/A+AS/94/73 Ramsbottom C.A.: J/A+A/475/765 J/A+AS/125/543 Ramsey L.W.: J/ApJ/554/742 J/ApJS/123/283 Rand R.J.: J/AJ/132/242 Randall S.W.: J/ApJ/600/729 J/ApJ/636/200

Randich S.: J/A+A/298/115 J/A+A/325/647 J/A+A/372/862 J/A+A/448/189 J/A+A/452/245

J/A+A/465/185 J/A+A/475/539 J/AJ/110/1229 Randich S: J/AN/319/201 Randich S.: J/AN/319/215 RANDONE I.: J/A+AS/107/285 Randone I.: J/A+AS/120/521 Rangel E.M.: J/A+A/431/565 ransom S.M.: J/MNRAS/359/1524 Ranta M.: J/AJ/133/1947 Rantakyro F.: J/A+A/471/103 Rao A.P.: J/A+AS/106/29 J/MNRAS/329/775 J/MNRAS/360/1305 Rao N.K.: J/A+A/353/287 Rao S.M.: J/ApJ/636/610 Rapacioli M.: J/A+A/469/575 Rapaport M.: I/272 I/303 J/A+A/320/652 J/A+A/383/1054 J/A+A/409/361 J/A+AS/125/399 Rappaport B.N.: VI/42 Raschotin S.I.: J/A+AS/132/233 Rashmi: J/A+AS/131/137 Raskin G.: J/A+A/449/305 J/A+A/463/243 Rassia E.: J/A+A/413/1037 Rastorguev A.S.: III/229 J/A+A/323/775 J/AZh/72/495 J/PAZh/23/90 J/PAZh/23/931 J/PAZh/24/443

J/PAZh/25/689 Ratag M.A.: J/A+AS/126/297 RATCLIFF S.J.: J/AJ/107/1453 Ratcliffe A.: J/MNRAS/300/417 J/MNRAS/306/592 Rathborne J.M.: J/ApJ/639/227 Ratnatunga K.: J/AJ/116/102 J/ApJS/121/417 RATNATUNGA K.U.: II/121 Ratnatunga K.U.: J/AJ/116/584 V/65 Ratorguev A.S.: J/PAZh/27/132 Rattenbury N.J.: J/MNRAS/330/137 Ratzlaff P.W.: J/ApJS/150/19 Rau A.: J/A+A/397/711 J/A+A/438/1175 Rauch M.: J/ApJ/596/768 Rauch T.: J/A+A/458/541 J/A+A/470/317 Rauer H.: J/A+A/365/110 J/A+A/378/116 J/A+A/379/564 Rauscher B.J.: J/AJ/109/1608 Rauterberg G.: J/A+A/390/39 Rauw G.: J/A+A/395/499 J/A+A/418/841 J/A+A/454/1047 J/A+A/463/981 J/A+A/477/593 J/MNRAS/372/661

J/MNRAS/375/145 Raux J.: J/A+A/437/789 Rave H.A.: J/ApJS/145/245 Ravikumar C.D.: J/A+A/465/1099 Ravindranath S.: J/ApJ/600/L99 J/ApJ/613/200 Rawat A.: J/A+A/435/507 J/A+A/465/1099 Rawley G.: J/ApJS/96/303 RAWLINGS S.: J/MNRAS/263/425 Rawlings S.: J/MNRAS/307/293 J/MNRAS/344/283 J/MNRAS/357/1231 J/MNRAS/372/1621 J/MNRAS/372/741

J/MNRAS/373/1531 J/MNRAS/375/1349 J/MNRAS/379/1343 Rawlins K.: J/other/APh/26.282 Rawson D.: J/ApJS/142/71 Rawson D.M.: J/ApJ/508/491 J/ApJ/521/155 Raychaudhury S.: J/A+AS/140/327 J/AJ/110/463 J/MNRAS/274/1107 J/MNRAS/339/652 Rayez J.C.: J/A+A/430/1139 Rayle K.: J/A+AS/133/197 Rayle K.E.: J/A+A/367/597 J/A+AS/136/379 Raymond J.C.: J/AJ/131/527 Raymond S.N.: J/AJ/128/426 Razzaque S.: J/other/APh/26.282 J/other/APh/26.367 R.E.: J/A+A/395/129 Reach W.: J/A+A/469/575 Reach W.T.: J/A+A/372/364 J/A+AS/127/1 J/ApJ/562/446 J/ApJ/643/965 Read A.M.: J/A+A/373/63 J/MNRAS/382/279 Read M.: VI/114 VI/115 VII/249 Read M.A.: J/MNRAS/362/689 Reader J.: VI/16 Readhead A.C.S.: J/A+A/472/763 J/AJ/130/23 J/ApJS/100/1 J/ApJS/81/83 J/ApJS/95/345 J/ApJS/98/1

J/ApJS/98/33 J/ApJS/99/297 J/MNRAS/354/485 VIII/72 Real Instituto y Observatorio de la Armada en San Fernando: I/147 I/170A I/189 I/205 I/213 I/256 I/304 REBEIROT E.: III/111 III/147 Rebeirot E.: J/A+AS/97/603 Rebeschini M.: J/AJ/121/327 Rebolo R.: J/A+A/398/363 J/A+AS/102/361 J/AN/325/705 J/MNRAS/341/1066 J/MNRAS/360/340 Reboul H.: J/A+AS/111/169 Rebull L.: J/A+A/468/353 J/AJ/127/2659 J/AJ/128/787 Rebull L.M.: J/AJ/119/3026 J/AJ/121/1676 J/AJ/123/1528 J/AJ/127/2228 Recillas E.: J/A+A/349/619 Recillas-Cruz E.: III/84B J/ApJ/453/616 Recio Blanco A.: J/A+A/440/901

Recio-Blanco A.: J/A+A/391/945 J/A+A/417/597 J/A+A/432/851 J/A+A/434/235 J/A+A/464/953

J/A+A/464/967 Recondo-Gonzalez M.C.: J/A+AS/121/461 Rector T.A.: J/AJ/130/923 Reddy B.E.: J/MNRAS/340/304 J/MNRAS/349/757 J/MNRAS/367/1329 J/MNRAS/369/1677 Reddy Naveen A.: J/ApJ/554/803 Redmann R.O.: II/41 Redondo I.: J/A+A/380/609 Reduzzi L.: J/A+A/330/423 J/A+AS/110/131 J/other/ApL/30.1 Ree C.H.: J/AJ/127/958 Reed B. Cameron: J/ApJS/115/271 REED B.C.: J/A+AS/77/447 J/AJ/106/1533 Reed B.C.: J/AJ/106/2291 J/PASP/105/1465 V/92A Reed C.: V/125 Reed C.B.: J/AJ/112/2855 Reed D.: J/MNRAS/351/265 Reed M.D.: J/A+A/459/557 Reegen P.: J/MNRAS/357/345 REES Jr. R.F.: J/AJ/106/1524 Rees N.: VIII/31 Rees R.F.: J/ApJ/443/124 Rees W.G.: J/MNRAS/229/589 Reese C.: J/A+A/430/1139 Reese D.F.: J/A+AS/106/333 Reffert S.: J/A+A/433/151 Refflinghaus F.: J/other/APh/26.282 Refregier A.: J/MNRAS/382/279 REGAN M.W.: J/AJ/105/499 Regan M.W.: J/ApJ/633/871 J/ApJS/146/353 Reglero V.: J/A+AS/119/271 J/A+AS/123/1 J/A+AS/123/59 Regnault N.: J/A+A/389/149 J/A+A/437/789 Rego M.: J/A+AS/118/7 J/A+AS/120/323 J/AJ/105/427 J/ApJS/105/343 J/ApJS/122/415 J/ApJS/95/387

Reich P.: J/A+A/323/739 J/A+A/327/1177 J/A+A/363/141 J/A+A/376/861 J/A+AS/106/303 J/A+AS/109/147 J/A+AS/126/413 J/A+AS/131/11 J/A+AS/75/67 J/A+AS/83/539 J/A+AS/85/805

Reich. P.: J/A+AS/91/337 REICH P.: J/ApJS/93/145

Reich W.: J/A+A/323/739 J/A+A/327/1177 J/A+A/363/141 J/A+A/376/861 J/A+A/448/411 J/A+AS/106/303 J/A+AS/109/147 J/A+AS/126/413 J/A+AS/131/11 J/A+AS/58/197 J/A+AS/75/67 J/A+AS/83/539 J/A+AS/85/805

Reichard T.A.: J/AJ/123/567 J/AJ/125/1711 J/ApJS/165/1 Reichart D.: J/AJ/130/968 Reichen M.: J/other/Ap+SS/163.275 Reichert G.A.: J/A+A/369/57

Reid I.N.: III/198 J/AJ/123/1570 J/AJ/123/2806 J/AJ/123/2822 J/AJ/123/2828 J/AJ/123/3356 J/AJ/125/3302 J/AJ/126/2421 J/AJ/126/2449 J/AJ/126/3007 J/AJ/127/2856 J/AJ/128/463 J/AJ/130/1871 J/AJ/132/891 J/AJ/133/2258 J/AJ/133/439 J/ApJ/564/421 J/MNRAS/301/1031 J/MNRAS/325/931 J/PASP/117/676

Reid: J/PASP/109/849 Reid M.: J/AJ/128/712 J/MNRAS/330/137 Reid M.A.: J/ApJ/625/891 J/ApJ/644/990 Reid M.J.: J/ApJ/393/149 J/ApJ/615/702 J/ApJS/160/220 J/ApJS/166/526 REID N.: J/AJ/103/514 Reid N.: J/ApJ/348/98 J/MNRAS/257/257 Reid R. I.: J/ApJS/124/285 Reid W.: J/MNRAS/335/1085 J/MNRAS/355/44 Reid W.A.: J/MNRAS/365/401 J/MNRAS/373/521 Reif K.: J/A+AS/145/223 J/A+AS/58/197 J/A+AS/85/805 Reig P.: J/A+A/451/835 J/A+A/462/1081 J/A+AS/119/271 J/A+AS/123/1 Reijns R.A.: J/A+A/360/472 J/A+A/445/503 Reimann: J/A+AS/120/429 Reimer O.: J/ApJ/615/897 J/ApJS/123/79 J/ApJS/135/155 REIMERS D.: J/A+A/294/377

Reimers D.: J/A+A/358/77 J/A+A/363/1040 J/A+A/375/366 J/A+A/378/556 J/A+A/391/L11 J/A+A/406/535 J/A+A/430/223 J/A+A/431/143 J/A+A/439/129 J/A+A/458/427 REIMERS D.: J/A+AS/110/469 Reimers D.: J/A+AS/115/235 J/A+AS/127/145 J/A+AS/128/507 J/A+AS/134/483 Reindl: J/A+A/404/423 Reinecke E-M.: J/A+A/378/556 Reiners A.: J/A+A/398/647 J/A+A/412/813 J/A+A/446/267 J/AJ/132/663 Reinfrank R.F.: J/MNRAS/371/898 REINHARDT M.: VII/6 Reinsch K.: J/A+A/318/111 J/A+A/335/467 J/A+A/347/47 J/A+AS/122/201 J/A+AS/132/341 Reiprich T.H.: J/ApJ/626/L5 J/ApJS/129/435 REIPURTH B.: J/A+AS/104/233 Reipurth B.: J/AJ/127/1117 J/ApJ/575/354 V/104 Reis Neto: J/A+A/384/650 Reisenegger A.: J/A+A/445/819 Reiss D.J.: J/A+A/415/863 REISWIG D.E.: J/AJ/106/2429 Reitzel D.B.: J/AJ/124/234 Reiz A.: J/A+AS/130/133 J/A+AS/53/181 Rejkuba M.: J/A+A/323/113 J/A+A/369/812 J/A+A/406/75 J/A+A/423/97 J/A+A/438/521 J/A+A/448/983

J/A+A/466/157 J/AJ/127/2114 J/AJ/127/24 J/ApJ/600/716 Rekola R.: J/A+A/427/769 J/A+AS/132/305 J/A+AS/143/357 Relano M.: J/A+A/430/911 J/A+A/466/905 Remillard R.: J/A+A/407/1039 Remy M.: J/A+A/340/371 Renan de Medeiros J.: J/A+A/451/1053 Renaud M.: J/ApJ/636/765 J/ApJS/170/175 Renault C.: J/A+A/303/137 RENAULT C.: J/A+AS/109/447 Rengelink R.: J/A+A/452/119 J/AJ/123/1784 J/MNRAS/339/913 J/MNRAS/346/627 J/MNRAS/366/1265 Rengelink R.B.: J/A+A/339/34 J/A+AS/124/259 Rengstorf A.: J/AJ/127/1158 Rengstorf A.W.: II/266 J/AJ/131/1923 Renson P.: III/152 III/162A III/199A J/A+A/378/113 J/A+AS/121/57 J/A+AS/127/201 Renzini A.: J/A+A/378/556 J/A+A/391/945 J/A+A/392/395 J/A+A/417/597 J/A+A/434/53 J/A+A/435/657

J/A+A/437/883 J/A+A/454/423 J/A+A/465/799 J/A+A/478/83 RENZINI A.: J/A+AS/111/17 Renzini A.: J/AJ/127/2738 J/AJ/130/1693 Renzini B.: J/A+A/331/70 Requena-Torres M.A.: J/A+A/455/971 Requieme Y.: I/196 I/272 I/303 REQUIEME Y.: I/63 Requieme Y.: J/A+AS/136/531 J/A+AS/89/311 Resch G.: J/PASJ/52/997 Resconi E.: J/other/APh/26.282 Reshetnikov V.: J/A+A/382/513 Reshetnikov V.P.: J/PAZh/25/447 Ressler M.E.: J/ApJ/630/381 Rest A.: J/AJ/130/2453 J/AJ/131/1648 Retter A.: J/A+A/448/1107 J/other/NewA/12.446 Rettura A.: J/A+A/434/53 J/A+A/454/423 J/A+A/478/83 Retzlaff J.: J/ApJS/129/435 Reuland M.: J/ApJ/619/755 Reuter H.P.: J/A+AS/102/611 Reuter H.-P.: J/A+AS/122/271 VIII/43 Revnivtsev M.: J/A+A/418/927 J/A+A/475/775 Revnivtsev M.G.: J/A+A/423/469 J/PAZh/28/323 J/PAZh/30/430 J/PAZh/30/589 Rey S.-C.: J/AJ/116/1775 J/AJ/119/1824 J/AJ/122/3219 J/AJ/126/2840 J/AJ/127/958 Reyle C.: J/A+A/390/491 J/A+A/421/643 J/A+A/441/653 J/A+A/453/635 J/MNRAS/373/705 Reyniers K.: J/A+A/463/243 Reyniers M.: J/A+A/417/269 J/A+A/463/243 J/A+A/465/965 Reynolds A.P: J/A+AS/109/9 Reynolds A.P.: J/A+AS/134/287 Reynolds C.S.: J/AJ/121/337 Reynolds J.: J/AJ/115/1253

Reynolds J.E.: J/AJ/110/880 J/AJ/127/3609 J/AJ/130/2529 J/MNRAS/268/602 J/MNRAS/301/1019

J/MNRAS/325/1411 J/other/PASAu/10.310 J/PASJ/52/997 Reynolds R.J.: II/249 R.F.: J/MNRAS/326/959 R.G.: J/A+A/421/937 Rhee G.: J/A+A/310/31 J/A+A/310/8 J/A+AS/134/401 J/ApJ/416/36 J/ApJS/107/175 J/ApJS/120/147 Rhee J.: J/A+A/437/61 J/A+A/445/69 J/A+A/454/717 J/AJ/128/245 J/AJ/130/2804 Rhee M.-H.: J/A+AS/115/407 Rho J.: J/ApJ/562/446 J/ApJ/607/904 J/ApJ/643/965 Rhoads J.E.: J/AJ/127/213 Rhode K.L.: J/A+A/448/155 Rhode W.: J/A+A/390/39 J/A+A/406/L9 J/A+A/410/813 J/other/APh/26.282 Rhodes J.: J/ApJS/146/209 J/ApJS/159/41 J/MNRAS/356/568 Rhodes J.D.: J/A+A/364/712 Ribas I.: J/A+A/326/1044 J/A+A/333/471 J/A+A/398/239 J/A+A/426/247 J/A+A/450/735 J/A+A/459/321

J/A+A/478/507 Ribas S.J.: J/A+A/425/489 Ribeiro A.L.B.: J/ApJS/110/1 Ribeiro J.P.: J/A+A/458/997 RIBONI E.: J/A+AS/108/55 Ribordy M.: J/other/APh/26.282 Ricci R.: J/A+A/415/549 J/A+A/445/465 J/MNRAS/354/305 J/MNRAS/371/898 Riccobono J.: J/ApJS/162/428 Rice J.: J/ApJ/459/278 Rice J.B: J/A+A/399/315 Rice R.H.: J/A+AS/136/111 Rice W.: II/156A RICE W.: VII/109 Rich J.: J/A+A/303/137 J/A+A/389/149 J/A+A/454/185 RICH J.: J/A+AS/109/447 Rich M.: J/AJ/111/1596 Rich R.M.: J/A+A/391/945 J/A+A/435/657 J/AJ/110/1649 J/AJ/110/1774 J/AJ/112/171 J/AJ/112/2408

J/AJ/124/1190 J/AJ/125/1598 J/AJ/127/2738 J/AJ/127/2771 J/AJ/127/660 J/AJ/130/1693 J/AJ/130/825 J/AJ/131/332 J/ApJ/612/447 J/ApJ/619/L95 J/ApJS/164/38

Richard C.: J/A+A/341/469 Richard J.: J/A+A/448/893 Richards A.M.S.: J/A+A/472/805 J/MNRAS/358/1159 Richards E.A.: J/A+A/472/805 J/AJ/123/2402 J/ApJ/533/611 J/MNRAS/358/1159 Richards G.T.: J/AJ/121/2308 J/AJ/121/31 J/AJ/122/503 J/AJ/123/567 J/AJ/125/1711 J/AJ/126/1720

J/AJ/126/2125 J/AJ/126/2209 J/AJ/126/2579 J/AJ/129/2542 J/AJ/130/873 J/AJ/131/1 J/AJ/131/49 J/ApJ/633/638 J/ApJS/123/377 J/ApJS/126/133 J/ApJS/135/227 J/ApJS/155/257 J/ApJS/165/1 VII/243

Richardson A.: J/AJ/122/2587 Richardson D.C.: III/185 Richardson G.A.: IX/20A Richardson M.: J/ApJ/613/200 RICHARDSON T.: II/161 Richberg K.: J/ApJ/538/29 Richer H.B.: J/A+A/390/961 RICHER H.B.: J/AJ/100/1811 J/AJ/103/460

Richer H.B.: J/AJ/109/2480 J/AJ/109/650 J/AJ/109/663 J/AJ/112/491 J/AJ/114/1488 J/AJ/117/247

J/AJ/127/2771 J/AJ/129/729 J/ApJS/120/265 Richer J.S.: J/A+A/440/151 J/A+A/468/1009 J/A+A/472/187 J/MNRAS/379/663 Richer M.G.: J/A+A/336/667 J/A+A/466/75 Richichi A.: J/A+A/386/492 J/A+A/413/1037 J/A+A/431/773 J/A+A/434/1201 J/AJ/112/2786 Richmond M.: J/ApJS/123/377 Richmond M.W.: II/230 II/271A RICHMOND M.W.: J/AJ/107/1022 Richmond M.W.: J/AJ/121/2308 J/AJ/126/2209 J/AJ/126/2579 J/PASP/110/553 Richstone D.: J/AJ/129/2138 Richstone D.O.: J/ApJ/470/724 Richter G.: J/A+A/359/865 RICHTER G.: J/A+AS/105/433 Richter G.: J/A+AS/117/39 J/A+AS/131/295 J/A+AS/135/511 J/A+AS/137/299 Richter G.A.: J/A+A/310/384 J/A+A/361/1005 Richter G.M.: J/A+A/296/643 Richter O.-G.: VIII/68 Richter P.: J/ApJS/146/1 J/ApJS/146/125 J/ApJS/155/351 Richter S.: J/other/APh/26.282 Richtler T.: J/A+A/304/347 J/A+A/415/123 J/A+A/433/43 J/A+A/459/391 RICHTLER T.: J/A+AS/109/1 J/A+AS/112/367

Richtler T.: J/A+AS/127/167 J/A+AS/134/59 J/A+AS/134/75 J/A+AS/135/391 J/AJ/115/592

J/AJ/127/2114 J/MNRAS/367/156 Ricker G.: J/other/GCN/1070.1 Ricker G.R.: J/ApJ/589/225 J/ApJ/613/1179 Rickett B.J.: J/ApJS/165/439 Ricketts M.J.: J/MNRAS/274/1165 VII/19 Rickman H.: J/A+AS/146/73 Riddle D.: J/A+A/447/921 Riddle R.L.: J/ApJ/525/420 Rider C.J.: J/AJ/127/2210 Ridge N.A.: J/MNRAS/379/663 Ridgway S.T.: J/A+A/453/155 Rieger E.: J/ApJS/120/409 Riehokainen A.: J/A+AS/143/357 Rieke G.: J/AJ/131/1574 Rieke G.H.: J/AJ/132/231 J/ApJ/540/1016 J/ApJ/593/1093 J/ApJ/620/1010 J/ApJ/633/871 J/ApJ/636/1098

J/ApJ/640/167 J/ApJ/642/126 J/ApJS/151/387 J/ApJS/154/155 J/MNRAS/380/986 Rieke M.: J/ApJ/641/140 J/ApJS/161/9 Rieke M.J.: J/AJ/117/17 J/AJ/130/1 J/ApJ/540/1016 J/ApJ/633/871 J/ApJS/107/281 J/ApJS/149/327

J/ApJS/151/387 J/ApJS/154/155 J/ApJS/96/117 Riello M.: J/A+A/391/945 J/A+A/432/851 Riepe B.Y.: J/AJ/105/1571 Ries P.: J/AJ/127/539 Riess A.: J/AJ/130/1 J/ApJ/600/L99 Riess A.G.: J/AJ/130/2453 J/AJ/131/1648 J/ApJ/594/1 J/ApJ/602/571 J/ApJ/607/665 J/ApJ/613/200

J/ApJ/627/579 Rifatto A.: J/A+A/428/339 J/A+AS/109/341 J/A+AS/114/527 Riffeser A.: J/A+A/445/423 J/A+A/465/375 J/ApJ/635/290 Rigaut F.: J/A+A/323/139 J/AJ/132/271 Rigby J.: J/ApJS/154/155 Rigby J.R.: J/ApJ/640/167 J/ApJ/642/126 Righetti G.: J/A+AS/74/315 Rigler M.: J/ApJ/471/694 J/ApJS/102/289

Rigopoulou D.: J/A+A/403/493 J/A+A/434/149 J/A+A/440/61 J/ApJ/640/167 J/ApJS/154/155

J/MNRAS/351/1290 RILEY J.M.: J/A+AS/110/419 Riley J.M.: J/MNRAS/204/151 J/MNRAS/246/110 J/MNRAS/274/324 J/MNRAS/276/614 J/MNRAS/282/779

J/MNRAS/298/637 J/MNRAS/301/203 J/MNRAS/307/293 J/MNRAS/315/839 J/MNRAS/354/485 J/MNRAS/376/1251 J/MNRAS/379/1442 VIII/32

Riley P.A.: J/MNRAS/201/495 Rincon F.: J/A+A/475/1053 Rines K.: J/AJ/120/2338 J/AJ/124/1266 J/AJ/126/2152 J/AJ/128/1078 J/AJ/131/527 Ringwald F.A.: J/A+A/384/504 J/AJ/109/2368 Rios L.A.: J/AJ/118/1131 Riou A.: J/A+AS/125/399

Ripepi V.: J/A+A/318/797 J/A+A/351/526 J/A+A/399/271 J/A+A/403/857 J/A+A/428/1043 J/A+A/428/339 J/A+A/439/413 J/A+A/457/79 J/A+A/462/599 J/MNRAS/363/734 J/MNRAS/371/1503

Ripken J.: J/A+A/406/L9 J/A+A/410/813 Risaliti G.: J/A+A/451/457 J/ApJS/121/473 Risi A.: J/A+AS/109/201 Ritchie C.: J/AJ/104/495 Rite C.: J/A+A/379/436 J/A+A/384/81 J/A+A/392/741 J/A+A/452/119 J/AJ/116/1 J/AJ/126/2268 Ritter H.: B/cb J/A+AS/123/273 V/110 V/113D V/123A RITTER H.: V/37 V/59 V/82 Ritter H.: V/99 Ritter L.A.: J/AJ/117/1313 RITZMANN B.M.: VI/53 Rivera E.J.: J/ApJ/634/625 Rivinius T.: J/A+A/389/931 Rivnius T.: J/A+A/459/137 Rix H.: J/ApJS/107/215

Rix H.-W.: II/253 J/A+A/377/442 J/AJ/122/2205 J/AJ/125/1107 J/AJ/127/105 J/AJ/131/1891

J/ApJ/598/260 J/ApJ/604/521 J/ApJ/619/193 J/ApJ/633/844 Rizvanov N.: J/A+A/375/670 J/A+A/444/625 Rizvi A.: J/AJ/130/234 Rizza E.: J/AJ/126/119 Rizzi L.: J/A+A/431/1189 J/A+A/452/179 J/A+A/467/1025 Rizzi M. Villata N.: J/A+AS/130/109 Rizzo A.: J/other/APh/26.282 Rizzo D.: J/A+A/403/857 J/A+A/428/1043 J/A+A/457/79 R.L.: VIII/73 Robb R.M.: J/A+A/415/289 J/A+A/466/1031 Robberto M.: J/A+A/453/911 J/AJ/129/1534 Robbins S.: J/other/APh/26.282 Robert C.: J/ApJS/144/21 J/ApJS/99/173 Robert M.S.: J/ApJ/438/181 Roberts B.: IX/35 Roberts D.A.: J/AJ/104/704 ROBERTS G.: J/A+AS/93/383 J/MNRAS/267/711 Roberts G.: J/MNRAS/325/931 J/other/SAAOC/15.78 Roberts L.C.Jr: J/AJ/119/3084 Roberts M.S.: J/AJ/108/1350 J/AJ/112/601 J/AJ/113/1011 J/AJ/116/180 J/ApJ/405/94 ROBERTS M.S.: J/ApJ/413/604 J/ApJ/425/122 Roberts M.S.: J/ApJ/438/181 J/ApJ/446/622 J/ApJS/104/287 J/ApJS/98/595 Roberts M.S.E.: J/MNRAS/359/1524 Roberts S.: J/MNRAS/352/478 J/MNRAS/357/819 Roberts T.P.: J/A+A/476/1191 J/MNRAS/357/401 J/MNRAS/370/25 Roberts W.H.: J/AJ/106/1839 ROBERTS W.H.: J/AJ/106/591 ROBERTS W.J.: J/A+AS/109/293 Roberts W.J.: J/A+AS/109/313 Robertson J.G.: VIII/46 Robertson J.W.: J/ApJS/115/185 Robertson-tessi M.: J/ApJS/157/371 ROBICHON N.: J/A+A/299/696 Robichon N.: J/A+A/331/81 J/A+A/345/471 ROBIN A.: III/111 III/147 Robin A.: J/ApJ/552/289 Robin A.C.: J/A+A/265/32 J/A+A/390/491 J/A+A/421/643 J/A+A/453/635 J/MNRAS/373/705 Robinson A.: J/ApJS/115/185 Robinson G.: J/MNRAS/291/261 J/MNRAS/301/640 Robinson J.G.: J/AJ/130/1733 Robinson P.E.: J/AJ/126/2048 Robinson S.E.: J/ApJ/637/1102 J/ApJS/169/430 Robrade J.: J/A+A/406/L9 J/A+A/461/669 Robson E.I.: J/AJ/130/1418 Rocca-Volmerange B.: III/196 III/232 J/A+A/475/791 J/MNRAS/302/222 J/MNRAS/305/297 J/MNRAS/347/837

J/MNRAS/351/1290 ROCCA-VOLMERANGE B.: VI/47 VI/47 Rocha-Pinto H.J.: J/A+A/423/517 J/A+A/475/217 J/AJ/128/245 J/MNRAS/298/332 Roche N.D.: J/MNRAS/361/607 Roche P.F.: J/MNRAS/361/211 Rocher P.: B/comets J/A+A/311/710 J/A+A/329/329 ROCHOWICZ K.: J/A+AS/108/669 Rockosi C.M.: J/AJ/121/2308 J/AJ/121/2331 J/AJ/123/567 J/AJ/126/2579 J/ApJ/588/824 Rodgers A.: J/ApJ/542/257 J/ApJ/598/597 Rodgers A.W.: J/AJ/106/1839 RODGERS A.W.: J/AJ/106/591 Rodgers A.W.: J/AJ/114/326 J/AJ/127/334 VII/216

Rodighiero G.: J/A+A/364/517 J/A+A/427/23 J/A+A/427/773 J/A+A/434/149 J/AJ/129/2074

J/MNRAS/371/1891 Rodmann J.: J/PASP/118/1690 Rodon M.: J/A+A/410/671

Rodono M.: J/A+A/310/173 J/A+A/327/1114 J/A+A/364/205 J/A+A/367/910 J/A+A/400/659

J/A+AS/115/41 J/A+AS/125/11 J/A+AS/138/87 Rodrigues I.: J/A+AS/128/545

Rodriguez E.: J/A+A/328/235 J/A+A/336/920 J/A+A/340/196 J/A+A/359/597 J/A+A/372/588

J/A+A/407/1059 J/A+A/450/715 J/A+A/456/261 J/A+AS/100/571 RODRIGUEZ E.: J/A+AS/101/421 J/A+AS/106/21 Rodriguez E.: J/A+AS/144/469 RODRIGUEZ E.: J/A+AS/93/189 Rodriguez E.: J/MNRAS/353/310 Rodriguez J.: J/A+A/467/585 Rodriguez L.F.: J/ApJ/463/205 J/ApJS/125/427 Rodriguez M.: J/ApJS/120/113 J/MNRAS/362/301 J/MNRAS/368/253 Rodriguez M.L.: J/A+AS/109/201 Rodriguez-Caderot G.: J/ApJS/95/387 Rodriguez-Espinosa J.M.: J/A+AS/129/455 J/AJ/110/212 Rodriguez-Fernandez N.J.: J/A+A/455/971 Rodriguez-Franco A.: J/A+A/455/971 Rodriguez-gil P.: J/MNRAS/373/687 Rodriguez-Merino L.: J/ApJ/601/485 J/ApJ/613/312 Rodriguez-Pascual P.M.: J/A+A/369/57 J/A+AS/121/461 RODRIGUEZ-ULLOA J.A.: J/AJ/106/1547 Roeder U.K.: III/122 Roederer I.U.: J/ApJ/617/L41 Roehring A.: J/A+A/390/39 Roeland Rengelink: VIII/69A Roennback J.: J/A+A/335/85 Roeser H.-J.: J/A+A/285/785 ROESER H.-J.: J/A+A/297/61 Roeser H.-J.: J/A+AS/118/65 ROESER S.: I/132 Roeser S.: I/146 I/154 ROESER S.: I/155 I/159A Roeser S.: I/193 ROESER S.: I/194 I/206 Roeser S.: I/208 Roeser S: I/210 Roeser S.: I/250 I/253 ROESER S.: III/155

Roeser S.: III/182 III/254 J/A+A/333/619 J/A+A/438/1163 J/A+A/440/403 J/A+A/468/151

J/A+A/477/165 J/AN/325/740 J/AN/328/889 ROESSIGER S.: VI/52 VI/58 ROEssiger S.: VI/68 Roettgering H.: J/AJ/131/1891 Roettgering H.J.A.: J/A+A/402/171 J/A+A/406/593 J/A+A/456/791 J/A+A/461/823 J/A+A/471/1105

J/A+AS/108/79 ROETTGERING H.J.A: J/A+AS/110/419 Roettgering H.J.A.: J/MNRAS/366/1265 Roger R.S.: J/A+AS/115/345 J/A+AS/65/485 J/A+AS/77/31 J/AJ/74/366 Rognvaldsson O.E.: J/ApJ/416/276 Rogoziecki P.: J/AJ/130/767 Roh D.-G.: J/ApJS/165/360 Rohlfs K.: III/15 III/76A Rohrback G.: II/156A Rohrmann R.: J/A+A/465/249 Rohrmann R.D.: J/A+A/385/986 Rojo P.M.: J/ApJS/133/119 Rola C.: J/ApJ/481/49 Roland J.: J/A+AS/134/401 J/ApJS/107/175 Roll J.: J/AJ/129/363 Rolland A.: J/A+A/328/235 J/A+AS/100/571 ROLLAND A.: J/A+AS/101/421 J/A+AS/106/21 Rolland A.: J/A+AS/129/9 ROLLAND A.: J/A+AS/93/189 Rolland A.: J/MNRAS/353/310 Rolland L.: J/A+AS/137/451

Rolleston W.R.J.: J/A+A/315/95 J/A+A/347/69 J/A+A/367/86 J/A+A/419/1123 J/A+A/436/687

J/A+AS/126/357 J/A+AS/126/81 J/A+AS/143/409 Rollinde E.: J/A+A/376/28 Roman N.G.: IV/12 IV/18 V/55 VI/42 Romanenko L.G.: I/292 Romani: J/A+A/386/1001 Romani R.W.: J/ApJ/590/109 J/ApJ/609/564 J/ApJ/626/95 Romaniello M.: J/A+A/397/473 J/A+A/418/525 J/A+A/461/397 J/AJ/121/2020 J/ApJS/155/271 Romanishin W.: J/ApJ/626/L5 Roman-Lopes A.: J/A+A/461/949 J/A+A/471/813 Roman-lopes A.: J/AJ/128/2364 Roman-Lopes A.: J/AJ/131/951 Romano D.: J/A+A/352/117 Romano P.: J/A+A/448/L9 J/A+A/474/793 Romanowsky A.J.: J/A+A/448/155 J/MNRAS/369/120 J/MNRAS/369/97 Romanyuk I.I.: J/other/Ap/46.234 Romer A.K.: J/ApJS/126/209 J/ApJS/140/239 J/MNRAS/341/1093 Romero G.E.: J/A+A/370/468 J/A+A/409/857 Romero-Colmenero E.: J/ApJ/644/829 J/MNRAS/311/456 Romney J.D.: J/PASJ/52/997 Roncin J.-Y.: J/A+AS/128/361 Rood H.J.: J/ApJS/125/35 VII/177 ROOD H.J.: VII/36 VII/56 Rood H.J.: VII/95 ROOD H.J.: VII/96 Rood R.T.: J/A+A/324/915 J/AJ/108/1350 J/ApJ/522/983 J/ApJS/165/338 VI/65 Ropbertson J.G.: VIII/70A VIII/78 VIII/81 Ros E.: J/AJ/130/2473 J/ApJ/609/539 Rosa M.R.: J/A+A/347/841 J/A+A/381/941 J/A+A/422/129 ROSADO M.: J/A+A/280/365 Rosado M.: J/AJ/124/3157 ROSAT Consortium: IX/30 ROSAT Scientific Team: IX/21 IX/28A IX/38 Rosati P.: J/A+A/393/425 J/A+A/424/545 J/A+A/432/381 J/A+A/434/53 J/A+A/451/457 J/A+A/454/423

J/A+A/478/83 J/AJ/123/2223 J/AJ/128/1990 J/ApJ/569/689 J/ApJ/595/685 J/ApJ/600/L99 J/ApJ/612/202 J/ApJ/613/200 J/ApJ/621/53 J/ApJ/644/30 J/ApJS/129/435 J/ApJS/139/369 J/ApJS/150/1 J/ApJS/155/271 J/ApJS/155/73 J/ApJS/161/21

Rose J.A.: J/AJ/123/1216 J/AJ/127/1513 J/AJ/130/2666 J/AJ/130/957 J/AJ/131/1280 J/AJ/131/866

J/ApJS/152/251 J/PASP/110/1434 Rose M.B.: J/AJ/132/393 Rose S.J.: J/A+A/416/383 J/A+A/424/363 J/A+A/432/1151 J/A+A/433/745 J/A+A/455/1157 Rosebery T.: J/AJ/122/2587 Roseboom I.: J/MNRAS/372/425 J/MNRAS/375/68 Roseboom I.G.: J/MNRAS/368/651 Rosen S.: J/A+A/474/473 J/A+A/476/1191 Rosen S.R.: J/A+A/469/27 Rosenberg A.: J/A+AS/144/5 J/A+AS/145/451 Rosenberg J.L.: J/AJ/109/2376 J/AJ/129/1311 J/AJ/130/2613 J/ApJS/130/177 Rosenbush A.E.: J/PAZh/26/535 Rosenzweig P.: J/AJ/127/1158 J/other/RMxAA/42.75 Rosino L.: J/A+A/338/1006 J/MNRAS/289/745 Roslund C.: J/other/IBVS/5633 Rosner R.: J/A+A/341/751 J/ApJ/315/687 J/ApJ/325/798 J/ApJ/348/253 J/ApJ/348/557 J/ApJ/351/492

J/ApJ/414/846 J/ApJS/102/75 Rosolowsky E.: J/ApJS/149/343 Ross N.: J/MNRAS/372/425 J/MNRAS/375/68 ROSSELLO G.: J/A+AS/94/519 Rossetti E.: J/A+A/413/343 Rossi C.: J/A+A/258/423 J/A+A/420/571 J/A+A/458/225 J/A+AS/114/341 J/A+AS/136/95 J/A+AS/92/685 Rossi E.: J/A+AS/103/541 ROSSI L.: J/ApJS/71/983 Rossi S.: J/A+A/431/143 J/A+A/439/129 J/AJ/112/1188 J/AJ/117/2329 J/AJ/117/981 J/AJ/119/2866

J/AJ/130/2804 J/ApJ/562/528 J/MNRAS/320/451 Rossinot P.: J/ApJ/638/293 Rosso C.: IX/10A J/A+A/325/1115 J/ApJS/117/319 ROSTOPCHINA A.: J/A+AS/112/457 Rostopchina A.N.: J/A+A/327/145 J/AZh/73/194 J/AZh/78/60 Rostopchina: J/A+A/448/1075 J/AZh/77/420 Roth J.: J/AJ/112/2408 Roth K.: J/AJ/127/2455 Roth M.: J/A+A/316/102 J/A+A/357/1020 J/A+A/416/475 J/A+AS/119/413 J/AJ/128/2239 J/MNRAS/367/513 Roth M.M.: J/A+A/306/9 Roth M.R.: J/AJ/125/1940 Rothermel H.: J/ApJS/94/551 Rothschild: J/A+A/395/129 Rothschild R.: II/77 J/A+A/342/736 Rott C.: J/other/APh/26.282 Rotter A.: J/A+A/459/875 Rottgering H.: J/A+A/362/519 J/A+A/444/365 J/AJ/122/2205 J/AJ/123/637 J/AJ/125/1107 J/ApJ/604/521

J/MNRAS/339/913 Rottgering H.J.A.: J/A+A/374/861 J/A+A/431/793 J/A+A/432/69 J/A+AS/124/259 J/A+AS/143/303

J/AJ/123/1784 J/ApJ/591/640 J/ApJS/150/417 J/MNRAS/346/627 J/MNRAS/355/20 J/MNRAS/382/279 Rouan D.: J/A+A/363/62 J/A+A/466/949 J/A+A/468/937 Roueff E.: J/A+A/445/361 J/A+A/449/855 J/A+AS/117/561 J/A+AS/141/297 ROUSSEAU J.: III/111 III/147 Rousseau J.: J/A+A/363/62 J/A+A/387/1 J/A+AS/117/467 VII/234 VII/237 Rousseau J.M.: I/230 J/A+AS/124/437 Rousseau M.: I/196 Roussel H.: J/A+A/361/429 J/ApJ/633/871 Rousselot P.: J/A+A/407/1139 Rovithis P.: J/A+A/327/1017 J/A+A/348/184 ROVITHIS P.: J/A+AS/105/47 Rovithis-Livaniou H.: J/A+A/327/1017 J/A+A/348/184 ROVITHIS-LIVANIOU H.: J/A+AS/105/47 Rowan-Robinson: J/MNRAS/296/839

Rowan-Robinson M.: J/A+A/359/865 J/A+A/372/364 J/A+A/467/565 J/A+A/467/73 J/AJ/127/3075 J/AJ/129/1198 J/AJ/129/2074 J/MNRAS/253/485 J/MNRAS/279/459 J/MNRAS/302/222 J/MNRAS/305/297 J/MNRAS/306/708 J/MNRAS/308/897 J/MNRAS/325/1173 J/MNRAS/343/293 J/MNRAS/344/887 J/MNRAS/351/1290 J/MNRAS/355/97 J/MNRAS/358/333 J/MNRAS/358/397 J/MNRAS/372/1621 VII/221

Rowe J.: J/AJ/122/2587 J/ApJ/542/257 Rowe J.F.: J/AJ/127/334 J/AJ/129/729 Rowell G.: J/A+A/390/39 J/A+A/406/L9 Rowell G.P.: J/A+A/410/813 Roy A.: VIII/50 VIII/50 ROY A.J.: I/25 Roy A.L.: J/MNRAS/301/1019 Roy J.R.: J/AJ/120/1306 Roy J-R.: J/ApJ/424/599 Roy R.: J/A+A/465/331 Roy S.: J/MNRAS/329/775 J/MNRAS/360/1305 Royal Greenwich Obs.: I/147 I/170A I/189 I/205 I/213 I/256 Royal Observatory: I/34 Royer F.: J/A+A/336/953 J/A+A/381/105 J/A+A/393/897 J/A+A/419/533 J/A+A/463/671 J/A+A/470/1009 Rozas: J/A+AS/135/145 Rozas M.: J/A+A/354/823 R.P.: J/A+A/395/753 Ruatti C.: J/A+AS/125/399 Ruban E.V.: III/201 J/PAZh/32/628 J/PAZh/32/672 Rubenstein E.P.: J/AJ/112/2408 Rubiera M.: V/68A Rubin D.S.: J/ApJ/631/85 Rubin V.C.: V/31A RUBIN V.C.: V/6 Rubino-martin J.A.: J/MNRAS/341/1066 Rubino-Martin J.A.: J/MNRAS/360/340 Rubio M.: J/A+A/276/25 J/A+A/471/103 J/AJ/125/1940 J/ApJS/162/1 J/MNRAS/366/739 J/PASJ/53/971 Rubio-Herrera E.: J/ApJ/641/L13 Rucinski S.M.: J/A+A/295/101 J/A+AS/114/1 J/AJ/121/254 J/AJ/126/237 J/AJ/130/767 J/AJ/131/407 Rud M.V.: J/A+A/409/169 Rudkin M.: J/A+A/459/557 Rudnick G.: J/A+A/427/397 J/A+A/444/365 J/AJ/122/2205 J/AJ/125/1107 J/AJ/131/1891 J/AJ/134/1103

J/ApJ/604/521 J/ApJ/630/206 J/MNRAS/371/1777 Rudnick L.: J/ApJ/441/307 J/ApJ/456/234 J/ApJ/613/343 Rudolph A.L.: J/ApJS/162/346 Rudolph D.: V/68A Rudy R.J.: J/A+A/459/875 J/A+AS/147/5 J/AJ/121/362 Ruelas-Mayorga A.: J/other/RMxAA/32.143 Ruelas-Mayorga R.A.: J/A+AS/93/61 J/A+AS/97/587 J/other/RMxAA/22.43 Rufener F.: II/169 II/72 RUIZ M. T.: J/AJ/106/2575 Ruiz M.T.: J/A+AS/119/413 J/AJ/112/2408 J/ApJ/595/247 J/ApJS/133/119 J/ApJS/133/413 J/ApJS/162/1

J/MNRAS/362/301 J/MNRAS/368/253 Ruiz-Lapuente P.: J/A+A/437/789 Ruiz-Moreno M.: J/A+AS/122/271 RULAND F.: J/A+AS/42/391 Rumstay K.S.: J/ApJS/115/185 Runacres M.C.: J/A+A/395/499 Ruoff J.: J/A+A/366/565 Ruoppo A.: J/MNRAS/371/1503 Rupen M.P.: J/ApJ/636/1002 Ruprecht J.: VII/101A RUPRECHT J.: VII/31B VII/44B Rush B.: J/ApJ/453/616 J/ApJS/125/349 VII/157 Rusholme B.: J/MNRAS/341/1066 Rusin D.: VIII/72

Russeil D.: J/A+A/397/133 J/A+A/417/107 J/A+A/470/161 J/MNRAS/362/689 J/other/Ap/44.328

J/other/Ap/44.88 J/other/Ap/45.322 J/other/Ap/46.475 V/127 Russel J.N.: I/220 I/254 I/255 Russell A.P.G.: J/A+AS/106/165 RUSSELL J.L.: I/151 Russell J.L.: I/197A J/AJ/110/880 Russell K.S.: J/MNRAS/362/689 Russell S.C.: J/A+A/326/1069 Russeva T.: J/AJ/119/2274 Russo G.: J/A+A/318/797 Russo T.A.: J/ApJ/610/247 Rutily B.: J/A+A/332/L53 J/A+A/342/773 J/A+A/344/263 J/A+A/369/178 J/A+A/390/967 Rutledge D.: J/other/APh/26.282 Rutledge G.A.: J/PASP/109/883 Rutledge R.E.: J/ApJ/578/405 Rutscheidt E.: V/31A Rutten R.G.M.: J/A+A/177/131 Ruymakers G.: J/A+A/367/865

Ruzdjak D.: J/A+A/350/566 J/A+A/387/580 J/A+A/408/611 J/A+A/415/289 J/A+A/416/669

J/A+A/455/1037 J/A+A/463/1061 J/A+A/463/233 J/A+A/464/263 Ruzic Z.: J/A+AS/125/75 Ruzicka F.: J/ApJS/117/319 R.W.: J/A+A/386/1001 Ryabchikova T.A.: J/A+A/458/569 J/PAZh/28/41 J/PAZh/31/437 Ryabinkov A.I.: J/A+A/412/707 Ryabtsev A.N.: J/PAZh/28/41 Ryan J.: J/A+AS/143/145 Ryan P.J.: J/A+AS/122/235 Ryan R.E.: J/ApJ/639/724 Ryan S.G.: J/A+AS/139/135 J/AJ/117/981 J/AJ/98/1693 J/ApJS/123/639 Ryans R.: J/A+A/467/785 J/MNRAS/372/1117 J/MNRAS/379/647 J/MNRAS/379/816

Ryans R.S.I.: J/A+A/419/1123 J/A+A/434/1125 J/A+A/436/687 J/A+A/437/467 J/A+A/457/265

J/A+A/466/277 J/A+A/471/625 Ryan-Weber E.: J/AJ/124/1954 Ryan-weber E.V.: J/AJ/128/16 Ryan-Weber E.V.: J/ApJS/165/307 Rybak J.: J/A+A/447/735 Rybka S.P.: I/285 J/A+AS/121/243 J/other/KFNT/23.102 Ryden B.S.: J/ApJ/517/650 Ryder S.D.: J/AJ/121/225 J/AJ/123/873 J/AJ/124/1954 J/AJ/124/690 J/AJ/128/16 J/MNRAS/337/641 Rykoff E.S.: J/AJ/127/2436 Ryl’kov V.P.: J/A+AS/118/105 RYS S.: J/A+A/281/15 Rys S.: VIII/53 Ryynaenen K.: J/A+A/383/197

S.: J/A+A/327/61 J/A+A/364/391 J/A+A/404/775 J/A+A/404/775 J/A+A/476/199 Saad S.M.: J/A+A/419/607 J/other/Ap+SS/203.53 Saar E.: J/A+A/459/L1 J/A+A/462/811 J/AN/327/365 Saarinen V.: J/A+A/440/409 Saavedra N.: J/AJ/112/2408 Saba V.W.: J/AJ/107/751 Sabas V.: J/A+AS/135/503 Sabater J.: J/A+A/436/443 J/A+A/449/937 J/A+A/462/507 J/A+A/470/505 J/A+A/472/121 Sabatini S.: J/A+A/430/411 J/MNRAS/352/478 J/MNRAS/357/819 Sabbey C.: J/MNRAS/359/930 Sabbi E.: J/A+A/464/939 Sabby J.A.: J/AJ/119/1389 J/AJ/128/1324 J/AJ/128/1340 J/AJ/128/1840 Sableviciute I.: J/PASJ/56/1025 Sabogal B.E.: J/A+A/478/659 J/MNRAS/361/1055 Sabogal-Martinez B.-E.: J/PASP/114/233 Saccavino S.: J/A+AS/116/1 Sacchi N.: J/A+A/457/501 Sachkov M.E.: III/229 J/PAZh/24/443 J/PAZh/28/653 Sacre J.J.: J/A+AS/125/399 Sacuto S.: J/A+A/465/469 Sadakane K.: J/A+A/420/673 J/A+A/459/423 J/PASJ/57/13 J/PASJ/57/27 Sadat R.: B/bax J/A+A/361/429 Sadler E.M.: J/A+AS/106/1 J/A+AS/120/257 J/AJ/110/1774 J/AJ/112/171 J/AJ/123/873 J/AJ/124/1954

J/AJ/124/690 J/AJ/128/16 J/AJ/89/53 J/MNRAS/329/227 J/MNRAS/337/641 J/MNRAS/347/837 J/MNRAS/352/1245 J/MNRAS/354/305 J/MNRAS/371/898 J/MNRAS/372/425 J/MNRAS/375/931 VIII/70A VIII/78 VIII/81

Sadzakov S.: J/other/BOBeo/155.3 Saesen S.: J/A+A/463/243 Saffer R.: J/A+A/317/689 Saffer R.A.: J/A+A/347/69 Safizadeh N.: J/MNRAS/358/795 Sagar: J/A+A/374/504 SAGAR R.: I/185 Sagar R.: J/A+A/381/65 J/A+A/402/113 J/A+A/455/303 SAGAR R.: J/A+AS/111/75

Sagar R.: J/A+AS/114/537 J/A+AS/122/9 J/A+AS/135/391 J/A+AS/145/365 J/MNRAS/250/683

J/MNRAS/299/1 J/MNRAS/327/23 J/MNRAS/337/133 J/MNRAS/358/1290 J/other/BASI/29.519 Sagar S.: J/A+AS/127/167 Saglia: J/A+A/395/753 Saglia R.: J/A+A/444/365 J/MNRAS/371/1777 Saglia R.P.: J/A+A/398/49 J/A+A/407/423 J/A+A/415/123 J/A+A/420/97 J/A+A/421/41 J/A+A/427/397

J/A+AS/120/257 J/A+AS/141/449 J/ApJ/635/290 J/ApJS/106/1 J/MNRAS/292/499 J/MNRAS/305/259 J/MNRAS/321/277 J/MNRAS/326/473

Saha A.: J/AJ/118/1684 J/AJ/123/3154 J/AJ/125/1261 J/ApJ/463/26 J/ApJ/507/655 J/ApJ/508/491 J/ApJ/512/48 J/ApJ/514/614 J/ApJ/521/155 J/ApJ/522/802 J/ApJ/523/540 J/ApJ/551/973 J/ApJS/154/623 J/PASP/117/37

SAHAI R.: J/A+AS/93/121 Sahakian V.: J/A+A/390/39 J/A+A/406/L9 J/A+A/410/813

Sahal-Brechot S.: J/A+A/400/791 J/A+A/423/397 J/A+A/441/391 J/A+AS/105/243 J/A+AS/105/245 J/A+AS/107/349 J/A+AS/109/551 J/A+AS/115/351 J/A+AS/117/127 J/A+AS/119/369 J/A+AS/119/529 J/A+AS/122/163 J/A+AS/122/533 J/A+AS/127/543 J/A+AS/128/359 J/A+AS/129/155 J/A+AS/130/539 J/A+AS/131/141 J/A+AS/131/143 J/A+AS/133/227 J/A+AS/140/191 J/A+AS/140/193

Sahi M.: IX/20A J/ApJS/154/585 Sahnow D.J.: J/ApJS/146/1 Sahu D.K.: J/MNRAS/372/1315 Sahu K.: J/ApJ/612/481 Sahu K.C.: I/279 J/AJ/119/486 Sainton G.: J/A+A/437/789 Saintonge A.: J/AJ/130/2613 J/ApJS/157/228 Saito H.: J/PASJ/51/791 J/PASJ/53/971 SAITO M.: J/A+A/299/347 Saito M.: J/A+AS/104/529 J/ApJS/112/245 J/ApJS/89/57 J/MNRAS/307/236 SAITO M.: VII/140 VII/140 Saito M.: VII/209A Saito S.: J/ApJS/117/427 Saito T.: J/MNRAS/330/137 Sakai K.: J/PASJ/52/631

Sakai S.: J/AJ/122/2893 J/ApJ/507/655 J/ApJ/508/491 J/ApJ/512/48 J/ApJ/514/614 J/ApJ/521/155

J/ApJ/523/540 J/ApJ/529/723 J/PASJ/52/631 Sakamoto S.: J/ApJS/100/125 Sakamoto T.: J/A+A/397/899 J/other/GCN/1070.1 Sakano M.: J/ApJ/518/656 Sakhibullin N.A.: J/AZh/81/209 Sakurai T.: J/A+A/468/663 VII/244A Sala G.: J/A+A/465/375 Sala M.: J/MNRAS/367/1170 Salabert D.: J/A+A/390/717 Saladyga M.: V/36B Salama A.: J/A+A/357/839 J/A+A/390/1033 J/AJ/112/241 Salaman M.: J/A+A/369/552 J/A+A/396/933 J/A+AS/135/499 J/A+AS/145/67 Salaris M.: J/A+A/432/851 J/A+A/463/579 Salas L.: J/A+A/349/619 J/A+A/361/660 J/A+A/399/1135 Salasnich B.: J/A+A/361/1023 SALGADO M.: J/A+A/291/155 J/A+AS/108/455 Salim S.: J/AJ/130/825 J/ApJ/582/1011 Salimbeni S.: J/A+A/449/951 Sallman M.: II/271A Salo H.: J/A+AS/141/103 J/MNRAS/355/1251 Salome P.: J/A+A/412/657 Saloranta J.: J/A+A/447/921 Saloranta P.-M.: J/AJ/133/1947 Salpeter E.E.: J/AJ/99/1071 J/ApJ/317/102 J/ApJS/103/331 Salter C.: J/PASJ/52/997 Salucci P.: J/A+A/342/671 J/ApJS/99/501 SALUKVADZE G.N.: I/134 I/135 I/136 Salvati M.: J/ApJ/510/637 J/ApJS/121/473 Salvo M.: J/MNRAS/369/1880 Salvo M.E.: J/AJ/130/2453 J/AJ/131/1648 SALZER J.J.: J/AJ/107/1453

Salzer J.J.: J/AJ/109/2376 J/AJ/113/1197 J/AJ/113/22 J/AJ/116/2805 J/AJ/117/1668 J/AJ/117/2039 J/AJ/121/66 J/AJ/123/1292 J/AJ/124/3465 J/AJ/127/1943 J/AJ/127/1959 J/AJ/130/2571 J/AJ/130/2584 J/AJ/130/496

Samain D.: J/A+AS/113/237 Sambruna R.: J/AJ/115/1253 Sambruna R.M.: J/A+A/414/825 J/A+A/433/1163 Samec R.G.: J/AJ/116/2549 Samec Ronald G.: J/AJ/113/401 Samodurov V.A.: J/AZh/77/21 Samorski M.: J/A+A/390/39 J/A+A/406/L9 J/A+A/410/813 Samotoev V.A.: J/AZh/81/209 Sams B.J.: J/ApJ/466/254 Samurovic S.: J/MNRAS/363/769 Samus N.: J/A+AS/109/487 SAMUS N.: J/A+AS/112/439 Samus N.: J/A+AS/119/191 J/A+AS/125/1 J/AJ/112/2004 J/AJ/112/2020 J/AJ/114/189 J/AJ/116/2415 Samus N.N.: B/gcvs Samus’ N.N.: II/139B II/140 Samus N.N.: II/172B II/195B II/205 II/214A II/219 II/250 III/229 J/A+A/397/997 J/A+AS/109/479

J/other/IBVS/5721 J/PAZh/20/693 J/PAZh/21/509 Samus’ N.N.: J/PAZh/21/596 Samus N.N.: J/PAZh/22/269 J/PAZh/22/764 J/PAZh/22/847 J/PAZh/23/454 J/PAZh/24/443 J/PAZh/28/201 Sana H.: J/A+A/454/1047 J/A+A/463/981 J/MNRAS/372/661 Sanahuja B.: III/232 J/A+AS/95/87 SANCHEZ ALMEIDA J.: J/A+AS/103/293 Sanchez Ge.: J/AJ/127/1158 Sanchez Gu.: J/AJ/127/1158 SANCHEZ M.: J/A+AS/110/351 Sanchez N.: J/other/RMxAA/38.129 Sanchez S.: J/A+AS/102/611 Sanchez S.F.: J/MNRAS/376/125 Sanchez-Blazquez P.: J/MNRAS/366/1230 J/MNRAS/371/703 J/MNRAS/374/664 Sanchez-Fernandez C.: J/A+A/407/791 Sanchez-Janssen R.: J/A+A/471/17 Sanchez-Saavedra M.L.: J/A+A/391/519 J/A+A/399/457 Sanchez-Sutil J.R.: J/A+A/452/739 Sancisi R.: J/A+A/370/765 J/A+A/442/137 Sand D.J.: J/ApJ/627/32 Sandage A.: III/144 III/145 J/A+A/404/423 J/AJ/90/1681 J/ApJ/522/802 J/ApJ/551/973 J/PASP/113/267

VII/180 VII/51 Sande C.B.: V/102 V/105 V/109 V/95 Sandell G.: J/A+AS/106/165 Sander H.-G.: J/other/APh/26.282 Sanders D.B.: J/A+A/413/843 J/AJ/110/1993 J/AJ/126/1607 J/AJ/127/3235 J/ApJ/522/113 J/ApJS/143/277

J/ApJS/143/315 J/ApJS/63/821 J/ApJS/73/359 VII/114 SANDERS W.T.: VII/43 SANDFORD S.A.: J/ApJS/86/713 Sandquist E.L.: J/AJ/126/2954 J/AJ/128/3019 J/ApJ/518/262 J/ApJ/608/243 J/ApJ/611/323 J/ApJ/628/729 Sandqvist A.: J/A+A/295/585 J/A+AS/115/439 Sandrelli S.: J/MNRAS/309/739 SANDULEAK N.: III/113 Sanduleak N.: III/180 III/247 III/43 SANDULEAK N.: III/94 Sanduleak N.: J/AJ/80/972 J/AJ/81/350 SANDULEAK N.: J/ApJS/66/309 Sanduleak N.: J/ApJS/66/387 SANDULEAK N.: J/ApJS/71/549 Sankarankutty S.: J/A+A/363/239 Sanner J.: J/A+A/349/448 J/A+A/357/471 J/A+A/369/511 Sano Y.: J/PASP/114/1382 Santacana G.: J/A+A/465/331 Santagata N.: J/A+AS/64/247 Santander-Garcia M.: J/A+A/465/481 Santander-Vela J.D.: J/A+A/462/507 Santantonio L.: J/A+A/439/487 Santiago B.: J/MNRAS/297/885

Santiago B.X.: J/A+A/455/453 J/A+A/458/113 J/A+A/467/1003 J/AJ/112/359 J/ApJS/107/1

J/MNRAS/338/253 Santolamazza P.: J/A+A/472/705 Santos: J/A+A/404/775 Santos J.F.C.Jr: J/A+A/440/111 J/AJ/116/723 J/MNRAS/335/233 Santos Jr. J.F.C.: III/219 Santos N. C.: J/A+A/466/743 Santos N.: J/A+A/476/1347

Santos N.C.: J/A+A/356/590 J/A+A/375/205 J/A+A/379/999 J/A+A/388/632 J/A+A/390/267 J/A+A/392/215 J/A+A/398/363 J/A+A/406/373 J/A+A/415/1153 J/A+A/415/391 J/A+A/426/695 J/A+A/438/251 J/A+A/449/723 J/A+A/458/873 J/A+A/458/997 J/A+A/469/783 J/A+A/470/231 J/A+A/474/293

Santos-Lleo M.: J/A+A/369/57 J/A+AS/121/461 Santos-Sanz P.: J/A+A/468/L13 Sanval N.B.: II/40 Sanvito G.: J/A+A/402/37 J/A+A/417/499 Sanz-forcada J.: J/A+A/416/281 Sanz-Forcada J.: J/A+A/445/673 J/A+A/446/501 J/A+A/461/977 J/A+AS/146/103 Sanzovo G.C.: J/A+AS/120/301 SAO Staff: I/131A

Saracco P.: J/A+A/375/1 J/A+A/392/395 J/A+A/437/883 J/A+AS/137/495 J/A+AS/145/111

J/AJ/120/2206 J/AJ/122/2190 Sarajedini A.: J/A+A/396/473 J/A+A/398/479 J/AJ/109/1086 J/AJ/109/1112 J/AJ/109/269 J/AJ/112/2408

J/AJ/113/264 J/AJ/116/2395 J/AJ/119/1760 J/AJ/119/1793 J/AJ/122/3219 J/AJ/124/601 J/AJ/128/224 J/ApJ/559/1060 J/ApJ/632/266 J/ApJS/93/161 J/MNRAS/358/1215 J/MNRAS/377/300 SARAJEDINI A.: J/PASP/106/404 Sarajedini V.: J/AJ/116/102 J/ApJS/159/41 Sarajedini V.L.: J/AJ/123/1292 J/AJ/124/3465 J/AJ/127/1943 J/AJ/130/2571 J/ApJ/571/136 J/ApJ/599/1006

J/ApJS/121/417 J/ApJS/142/1 J/ApJS/157/175 J/MNRAS/371/221 Sarazin C.L.: J/ApJ/599/218 J/ApJ/600/729 J/ApJ/617/262 J/ApJ/636/200 SARAZIN C.L.: VII/56 Sarcander M.: II/62 Sareyan J.P.: J/A+A/304/406 J/A+A/407/1059 Sareyan J.-P.: J/A+A/426/247 Sargent A.I.: J/ApJ/555/932 J/ApJ/638/293 Sargent W.L.: J/ApJS/98/477 Sargent W.L.W.: J/AJ/122/2810 J/AJ/126/539 J/ApJ/472/509 J/ApJ/554/742 J/ApJ/596/768 J/ApJS/112/315

J/PASP/106/646 Sargsyan L.A.: VI/116 Saripalli L.: J/AJ/130/896 Sarma M.B.K.: J/A+A/297/359 Sarneczky K.: J/A+A/374/712 J/A+A/375/285 J/A+A/376/561 J/A+A/410/879 J/other/P+SS/53.925 Sarro L.M.: J/A+A/475/1159 Sartori: J/A+A/438/769 Sartori M.J.: J/A+A/361/1143 J/A+A/404/913 J/A+AS/128/561 Sarzi M.: J/A+A/342/671 J/A+A/360/439 J/AJ/127/105 Sasaki M.: J/A+AS/143/391 J/A+AS/147/75 Sasaki S.S.: J/PASJ/57/881 Sasaki T.: J/ApJ/648/7 J/PASJ/57/165 J/PASJ/58/313 Sasao T.: J/PASJ/52/631 Sasseen T.: J/ApJS/111/143 J/MNRAS/342/802 J/MNRAS/362/1371 J/MNRAS/369/156 Sasseen T.P.: J/ApJS/96/461 Sasselov D.D.: J/AcA/48/455 J/AJ/115/1016 J/AJ/115/1894 J/AJ/117/2810 J/AJ/118/2211 J/AJ/118/346

J/AJ/121/2032 J/AJ/121/861 J/AJ/121/870 J/AJ/122/1383 J/AJ/122/2477 J/AJ/123/3460 J/AJ/126/175 J/AJ/127/861 J/AJ/129/2856

Sastry K.V.L.N.: J/ApJ/499/517 Sato B.: J/PASJ/57/109 J/PASJ/57/13 J/PASJ/57/27 Sato F.: VII/244A Sato G.: J/ApJ/635/214 Sato H.: J/MNRAS/330/137 Sato K.: J/PASJ/52/631 V/69

Sato S.: J/AJ/117/446 J/ApJ/629/288 J/MNRAS/353/705 J/MNRAS/354/980 J/MNRAS/368/534

J/PASJ/55/1079 SATO S.: VII/60 Sato Y.: J/A+A/365/409 J/A+A/413/843 J/MNRAS/379/1343 J/PASJ/57/165 Satyapal S.: J/A+A/414/825 Sault R.J.: J/A+A/415/549 J/A+A/445/465 J/AJ/130/1358 J/MNRAS/289/225 Saumon D.: J/other/Sci/292.698 Saunders E.S.: J/MNRAS/375/1220 Saunders R.: J/A+AS/108/79 SAUNDERS R.: J/MNRAS/263/425 Saunders R.D.E.: J/MNRAS/341/1066 J/MNRAS/342/915 J/MNRAS/360/340

Saunders W.: J/AJ/130/957 J/AJ/131/1280 J/MNRAS/253/485 J/MNRAS/279/459 J/MNRAS/308/897

VII/221 VII/249 Saurer W.: J/A+AS/110/269 J/A+AS/117/1 J/A+AS/126/247 J/A+AS/127/101 SAURER W.: J/A+AS/93/553 Saurer W.: J/AJ/107/2101 Sauty S.: J/A+A/411/381 Sauvage M.: J/A+A/406/867 J/ApJS/155/401 Sauvageot J.-L.: J/A+AS/124/283 Sauval A.J.: J/A+A/347/348 SAVAGE A. [1: J/A+AS/112/347 Savage A.: J/MNRAS/238/1171 J/other/PASAu/10.310 VIII/34 VIII/50 SAVAGE B.D.: II/114 Savage B.D.: J/AJ/116/941 J/ApJ/508/200 J/ApJS/136/631 J/ApJS/146/1 J/ApJS/146/125 J/ApJS/155/351

J/ApJS/93/211 J/ApJS/97/1 Savage R.: J/MNRAS/341/1066 Savage R.S.: J/MNRAS/372/1621 Savaglio S.: J/A+A/318/347 J/AJ/119/486 J/AJ/127/2455 Savanov I.S.: J/AZh/72/855 J/AZh/72/864 J/AZh/73/862 J/AZh/75/586 J/AZh/77/96 Saviane I.: J/A+A/315/40 J/A+A/333/479 J/A+A/391/945 J/A+A/459/875 J/A+A/467/1025 J/A+AS/144/5

J/A+AS/145/451 J/AJ/127/660 Savine C.: J/A+A/451/1159 Savolainen T.: J/A+A/427/769 Sawada-Satoh S.: J/PASJ/52/631 Sawicki M.: J/AJ/121/1840 J/AJ/126/1208 J/ApJ/565/208 J/ApJ/627/621 J/ApJS/129/475 Sawyer Hogg H.: V/97 V/97 Saxe D.H.: J/AJ/123/567 J/AJ/126/2579 VII/149 Saxner M.: J/A+A/151/372 Sayagac A.T.: J/PAZh/26/843 Sayers J.: J/ApJ/638/293 Sazonov S.: J/A+A/418/927 J/A+A/474/793 J/A+A/475/775 Sazonov S.Y.: J/A+AS/129/1 Sazonov S.Yu.: J/A+A/423/469 Sbarufatti B.: J/A+A/469/503 J/A+A/474/793 Sbordone L.: J/A+A/414/503 J/A+A/464/201 J/A+A/465/815 Scalo J.: J/A+A/423/517 SCALTRITI F.: J/A+AS/102/343 Scaltriti F.: J/A+AS/121/119 Scannapieco E.: J/A+A/432/381 Scappini F.: J/MNRAS/276/57 Scaramella R.: J/A+A/337/17 J/A+A/342/1 J/A+A/379/21 J/A+A/379/35 J/A+A/403/857 J/A+A/428/1043

J/A+A/428/339 J/A+A/457/79 J/A+AS/130/323 J/MNRAS/296/599 J/MNRAS/312/540 Scardia M.: J/A+AS/134/545 J/ApJS/131/561 J/MNRAS/367/1170 Scarfe C.D.: J/AJ/129/1001 Scarpa R.: J/A+A/462/L9 Scatarzi A.: J/MNRAS/349/793 Scelsi L.: J/A+A/468/353 J/A+A/468/379 Schachter J.: J/A+A/355/651 J/ApJ/481/95 Schade D.: J/AJ/128/1558 J/AJ/96/1632 J/ApJ/464/79 J/ApJ/481/49 J/ApJ/499/112 J/ApJS/102/289

J/ApJS/153/397 J/ApJS/157/228 SCHADE D.: VII/143 Schaefer B.: J/AJ/127/1158 Schaefer B.E.: J/A+A/459/875 Schaefer J.: J/AJ/119/1901 Schaeffer A.J.: J/AJ/131/1501 Schaerer D.: III/232 J/A+A/381/941 J/A+A/422/129 J/A+A/441/981 SCHAERER D.: J/A+AS/101/415 J/A+AS/102/339 J/A+AS/103/97 Schaerer D.: J/A+AS/115/339 J/A+AS/128/471 J/A+AS/135/405 J/A+AS/136/35 SCHAERER D.: J/A+AS/96/269 J/A+AS/98/523 Schaerer D.: VI/102 VI/109 Schaffer M.A.: J/ApJ/631/85 Schahmaneche K.: J/A+A/437/789 Schakht N.A.: I/295 Schalander P.: J/other/IBVS/5633 SCHALLER G.: J/A+AS/101/415 J/A+AS/102/339 J/A+AS/103/97 J/A+AS/96/269 Scharf C.A.: J/ApJ/569/689 J/ApJS/140/265 Schartel N.: J/A+A/307/33 J/A+A/476/1191 J/ApJ/450/51 Schaye J.: J/ApJ/596/768 J/ApJS/163/80

Schechter P.L.: II/162 J/AJ/112/772 J/AJ/115/635 J/AJ/116/469 J/AJ/83/1549 J/AJ/88/1285

J/ApJ/314/493 J/ApJS/130/237 VII/203 Scheck M.: J/A+AS/142/275 V/76 Scheidegger R.: J/A+A/459/679 Schenner H.: J/AJ/127/1158 Schertl D.: J/A+A/385/87 J/A+A/422/627 J/A+A/448/1075 Schiavon R.: VI/120 Schiavon R.P.: J/A+AS/128/561 J/AJ/127/1513 J/AJ/130/2666 Schielicke R.: III/95 Schilbach E.: III/254 SCHILBACH E.: J/A+A/299/696

Schilbach E.: J/A+A/332/575 J/A+A/384/145 J/A+A/384/925 J/A+A/438/1163 J/A+A/440/403 J/A+A/468/151 J/A+A/477/165 J/A+AS/134/525 J/A+AS/137/305 J/AN/316/225 J/AN/325/740 J/AN/328/889 J/BaltA/9/646

Schild H.: J/A+A/439/265 Schild R.: IX/22 J/A+A/420/447 Schild R.E.: IX/15 J/AJ/131/527 J/ApJS/35/111 Schilizzi R.T.: J/A+A/402/171 J/A+A/435/863 J/PASJ/52/997

Schilke P.: J/A+A/321/293 J/A+A/466/1065 J/A+A/469/207 J/A+A/475/925 J/A+A/476/1243

J/ApJS/156/127 Schiller S.J.: J/AJ/109/359 J/PASP/106/1120 Schilling M.: J/A+A/390/39 J/A+A/406/L9 J/A+A/410/813 Schiminovich D.: J/AJ/130/825 J/ApJ/619/L95 J/ApJS/164/38 Schimpke T.: J/A+AS/93/463 Schindler S.: J/A+A/425/367 J/ApJ/514/148 Schinnerer E.: J/AJ/128/1974

Schirmer M.: J/A+A/379/436 J/A+A/379/740 J/A+A/384/81 J/A+A/441/905 J/A+A/443/103

J/A+A/458/245 J/A+A/462/875 Schlegel D.: J/AJ/122/503 J/AJ/123/3409 J/AJ/123/567 J/AJ/127/899 J/AJ/131/2018 J/ApJ/588/824

VII/185 Schlegel D.J.: J/AJ/121/2308 J/AJ/121/2331 J/AJ/124/1651 J/AJ/125/1817 J/AJ/126/1023 J/AJ/126/1720

J/AJ/126/2125 J/AJ/126/2579 J/AJ/127/3553 J/AJ/129/2542 J/AJ/131/1 J/ApJ/598/260 J/ApJ/607/426 J/ApJS/163/80 J/MNRAS/313/469 J/MNRAS/327/249 J/MNRAS/327/265

Schlegel E.: J/ApJ/644/829 Schlegel E.M.: J/ApJ/589/509 J/ApJS/150/19 J/ApJS/96/303 Schlenstedt S.: J/other/APh/26.282 Schlesinger F.: I/141 Schlickeiser R.: J/A+AS/131/11 Schlingman W.: J/AJ/131/2478 J/PASP/118/1690 Schloerb F.P.: J/ApJ/620/961 Schlosser W.: J/A+AS/127/1 Schmeja S.: J/A+A/413/1037 Schmele D.: J/A+A/390/39 Schmid H.M.: J/A+A/389/931 Schmider F.X.: J/A+A/390/717 Schmidt B.P.: J/A+A/415/863 J/AJ/128/712 J/AJ/130/2453 J/AJ/131/1648 J/ApJ/594/1 J/ApJ/602/571 SCHMIDT E.G.: J/AJ/106/2429 Schmidt E.G.: J/AJ/112/2769 J/AJ/128/1748 J/AJ/129/2007 J/AJ/130/832 SCHMIDT E.G.: J/ApJ/408/484 Schmidt E.G.: J/ApJS/104/101 SCHMIDT E.G.: J/ApJS/89/259 Schmidt E.G.: J/ApJS/96/605 J/PASP/108/1105 Schmidt G.: J/A+A/349/619 J/AJ/126/1023 J/ApJ/607/426 Schmidt G.D.: J/AJ/131/1674 J/ApJ/512/125 J/ApJ/569/23 SCHMIDT K.-H.: VII/161

Schmidt M.: II/207 J/A+A/329/482 J/A+A/371/833 J/A+A/393/425 J/AJ/107/1245 J/AJ/108/1163

J/AJ/130/873 J/ApJ/269/352 J/ApJ/616/1072 Schmidt R.W.: J/A+A/440/53 Schmidt S.J.: J/AJ/133/2258 J/AJ/133/439 Schmidt T.: J/A+A/468/1039 Schmidt-Kaler T.: J/A+A/325/693 J/A+AS/121/247 J/A+AS/99/591 J/MNRAS/289/729 J/MNRAS/306/279 Schmidt-Kaler Th.: J/AN/320/105 J/MNRAS/286/500 SCHMIDT-KALER TH.: VII/5A Schmidtke P.C.: J/AJ/114/2353 SCHMIDTKE P.C.: VII/143 Schmitt H.H.M.M.: J/A+AS/127/251 J/A+AS/132/155 J/A+AS/135/319 SCHMITT H.R.: J/A+A/283/805 Schmitt H.R.: J/AJ/117/238 J/ApJS/101/41 Schmitt J.: IX/29 J/MNRAS/274/1194 Schmitt: J/A+A/403/247 Schmitt J.H.M.M.: II/203A IX/10A IX/35 J/A+A/297/391 J/A+A/298/115 J/A+A/311/211 J/A+A/312/439

J/A+A/319/184 J/A+A/320/185 J/A+A/325/1115 J/A+A/341/163 Schmitt J.H.M.M: J/A+A/398/647

Schmitt J.H.M.M.: J/A+A/402/571 J/A+A/412/813 J/A+A/417/651 J/A+A/439/1137 J/A+A/461/669 J/A+A/475/677 J/A+AS/114/109 J/A+AS/118/481 J/A+AS/123/329 J/A+AS/128/29 J/ApJ/325/798 J/ApJ/348/557 J/ApJ/351/492 J/ApJ/448/683

SCHMITZ M.: II/136 II/184 Schmitz M.: II/209 II/216 II/225 II/71 II/98 SCHMITZ M.: IV/21 Schmitz M.: J/PASP/110/779 Schmoll J.: J/A+A/349/448 J/A+AS/124/157 J/ApJS/115/185 Schmutz W.: J/A+A/439/265 Schneid E.J.: J/ApJS/101/259 J/ApJS/107/227 J/ApJS/123/79 J/ApJS/94/551 Schneider D.: J/ApJS/148/243 J/other/APh/26.282 Schneider D. P.: J/ApJS/107/215 Schneider D.P.: J/A+A/371/833 SCHNEIDER D.P.: J/AJ/101/1561 Schneider D.P.: J/AJ/107/1245 J/AJ/107/1745 SCHNEIDER D.P.: J/AJ/108/1786 Schneider D.P.: J/AJ/113/669 J/AJ/116/1757 J/AJ/116/584 J/AJ/121/2308 J/AJ/121/2331 J/AJ/121/31

J/AJ/122/2810 J/AJ/122/503 J/AJ/123/3409 J/AJ/123/567 J/AJ/124/1651 J/AJ/124/2351 J/AJ/125/1711 J/AJ/125/1817 J/AJ/126/1023 J/AJ/126/1720 J/AJ/126/2125 J/AJ/126/2209 J/AJ/126/2579 J/AJ/126/539 J/AJ/127/1811 J/AJ/127/2838 J/AJ/127/3034 J/AJ/127/3553 J/AJ/127/899 J/AJ/129/2542 J/AJ/130/387 J/AJ/130/873 J/AJ/130/968 J/AJ/131/1 J/AJ/131/1674 J/AJ/131/2018 J/AJ/131/2722 J/AJ/131/49 J/AJ/131/84 J/AJ/132/676 J/ApJ/264/337 J/ApJ/409/28 J/ApJ/554/742 J/ApJ/588/824 J/ApJ/607/426 J/ApJ/633/638 J/ApJS/145/245 J/ApJS/155/257 J/ApJS/161/21 J/ApJS/163/80 J/ApJS/165/1 J/MNRAS/372/425 J/MNRAS/375/68 J/PASP/106/646 VII/149 VII/243

Schneider G.: J/AJ/117/17 J/AJ/130/1 J/ApJ/540/1016 J/PASP/110/1012 Schneider H.: III/59 Schneider J.: J/A+A/365/110 J/A+A/378/116 J/A+A/379/564 Schneider N.: J/A+A/473/149 J/A+A/476/1243 Schneider P.: J/A+A/441/905 J/A+A/444/365 J/A+A/462/875 J/A+A/470/821 Schneider S.: II/246 Schneider S.E.: J/A+A/405/99 J/A+A/408/465 J/A+A/408/67 J/AJ/129/1311 J/ApJ/459/606 J/ApJ/560/566

J/ApJS/119/159 J/ApJS/130/177 J/ApJS/81/5 Schnell A.: J/A+A/444/941 J/A+A/452/959 J/A+A/458/293 Schoedel R.: J/A+A/433/117 Schoeier F.L.: J/A+A/432/369 Schoenball M.: J/A+A/447/921 Schoenberner D.: J/A+A/339/123 J/A+AS/126/39 SCHOENEICH W.: II/200 Schoenfeld E.: I/119 Schoenfelder V.: J/A+AS/143/145 SCHOENING T.: J/A+AS/113/579 Schoenmaker T.: J/A+A/323/469 Schoenmakers A.P.: II/202 J/A+A/304/415 J/A+A/374/861 Schofield Jr. N.J.: IV/12 Schohn C.: V/100 V/84 Scholtz R.-D.: J/A+A/312/74 Scholz A.: J/A+A/416/555 J/A+A/417/557 J/A+A/419/249 J/A+A/429/1007 J/MNRAS/374/399 Scholz D.: J/A+AS/119/489 Scholz G.: J/A+A/304/235 J/A+A/367/236 Scholz M.: J/A+A/470/191 Scholz R.-D.: III/239 III/254 J/A+A/320/776 J/A+A/353/958 Scholz R.D.: J/A+A/384/145

Scholz R.-D.: J/A+A/390/491 J/A+A/392/851 J/A+A/438/1163 J/A+A/440/1061 J/A+A/440/403

J/A+A/442/211 J/A+A/468/151 J/A+A/477/165 SCHOLZ R.-D.: J/A+AS/104/161 Scholz R.-D.: J/A+AS/137/305 J/AN/316/225 J/AN/325/740 J/AN/328/889 Scholz R.D.: J/MNRAS/373/705 Schombert J.: J/A+A/403/937 J/AJ/124/3061 J/AJ/130/1002 SCHOMBERT J.M.: J/AJ/106/530 Schombert J.M.: J/AJ/109/618 J/ApJS/131/47 Schommer R.: J/ApJ/594/1 J/ApJS/102/29 Schommer R.A.: J/AJ/112/2408 J/AJ/127/2674 J/ApJ/562/303 J/MNRAS/300/665 Schordan P.: J/AN/321/277 Schrabback T.: J/A+A/441/905 Schraml J.: J/A+AS/35/23 Schramm T.: J/A+A/318/680 Schreiber W.: J/A+AS/144/123 Schreier E.: IX/3 J/ApJS/139/369 J/ApJS/155/271 Schreier E.J.: J/ApJ/595/685 Schreyer K.: J/A+AS/115/285 J/A+AS/124/385 J/AJ/129/1049 J/ApJS/161/361 Schrijver C.J.: J/A+A/325/1115 Schrijver H.: J/A+A/331/949 Schrnidtobreik L.: J/A+A/436/527 Schroder A.: J/AJ/124/1954 Schroder L.L.: J/AJ/119/727 Schroeder A.: J/A+A/363/62 J/A+AS/118/441 J/AJ/128/16 J/AJ/129/220 Schroeder A.C.: J/A+A/391/887 J/A+A/466/481 Schroeder C.: J/A+A/475/677 Schroeder F.: J/A+A/390/39 Schroeder J.: J/AJ/127/2838 J/ApJS/148/243 Schroeder S.E.: J/A+A/428/149 Schuberth Y.: J/A+A/433/43 J/A+A/459/391 Schuecker P.: J/A+A/420/61 J/A+A/425/367 J/ApJ/514/148 J/ApJS/129/435 J/ApJS/140/239 Schuehle U.: J/A+AS/126/281 Schuhle U.: J/A+A/375/591 Schuler S.C.: J/ApJ/602/L117 J/ApJ/632/L131

Schuller F.: II/243 J/A+A/362/199 J/A+A/392/971 J/A+A/400/903 J/A+A/403/141 J/A+A/405/531

J/A+A/446/551 J/A+A/453/535 J/MNRAS/381/1219 SCHULMAN E.: J/AJ/107/1022 SCHULTE-LADBECK R.E.: J/A+AS/111/527 Schultheis M.: II/243 J/A+A/338/581 J/A+A/349/L69 J/A+A/362/215 J/A+A/403/141 J/A+A/405/531

J/A+A/453/535 J/A+A/453/635 J/A+AS/123/115 J/A+AS/131/105 J/ApJ/552/289 J/MNRAS/337/519 J/MNRAS/373/705 J/MNRAS/381/1219

Schultz J.: J/A+A/350/513 Schulz B.: J/A+A/357/839 J/A+A/369/57 J/A+A/407/791 Schulz J.: J/A+A/392/1 J/A+A/398/89 Schulz N.S.: J/ApJ/549/441 J/ApJS/160/511 Schurch M.P.E.: J/A+A/461/631 Schurch N.: J/A+A/476/1191 Schuster H.E.: VI/30A SCHUSTER W.J.: II/180

Schuster W.J.: J/A+A/326/751 J/A+A/422/527 J/A+A/445/939 J/A+AS/117/317 J/A+AS/97/951

J/MNRAS/360/1345 J/MNRAS/371/1793 Schuurmans J.J.: J/A+A/472/705 Schwan H.: I/149A I/175 I/210 I/264 I/286 J/A+A/243/386 Schwarz H. E.: III/174 Schwarz H.E.: J/A+A/343/841 J/A+A/403/93 J/A+A/405/627 J/A+A/427/613 J/A+AS/119/37 Schwarz J.: IX/13 Schwarzenberg-Czerny A.: J/MNRAS/371/1405 Schweitzer A.E.: J/AJ/110/2747 Schweitzer E.: J/A+A/323/113 Schweizer F.: J/AJ/109/960 J/AJ/114/2381 Schweizer T.: J/A+A/410/813 Schwekendiek P.: I/207 Schwekendiek P: I/259 SCHWERING P.B.W.: II/181 Schwope A.: J/A+A/428/383 J/A+A/470/821 J/A+A/474/473 J/A+A/476/1191 Schwope A.D.: IX/32 J/A+A/335/467 J/A+A/347/47 J/A+A/466/41 J/MNRAS/311/456 Sciortino D.: J/A+A/461/977

Sciortino S.: J/A+A/341/751 J/A+A/345/521 J/A+A/355/651 J/A+A/397/147 J/A+A/399/919 J/A+A/404/927 J/A+A/417/945 J/A+A/421/175 J/A+A/430/287 J/A+A/430/941 J/A+A/450/993 J/A+A/455/903 J/A+A/459/199 J/A+A/460/133 J/A+A/462/245 J/A+A/463/275 J/A+A/464/211 J/A+A/474/495 J/ApJ/315/687 J/ApJ/325/798 J/ApJ/348/557 J/ApJ/351/492 J/ApJ/414/846 J/ApJ/582/382 J/ApJ/588/1009 J/ApJ/606/466 J/ApJ/608/781 J/ApJS/102/75 J/ApJS/160/319 J/ApJS/160/450 J/MNRAS/326/1161

Sclavi S.: J/A+A/451/435 Scoccimarro E.R.: J/A+AS/109/201

Scodeggio M.: J/A+A/361/863 J/A+A/403/857 J/A+A/409/439 J/A+A/428/1043 J/A+A/452/119 J/A+A/457/79 J/A+AS/137/75 J/A+AS/137/83 J/A+AS/142/65 J/A+AS/142/73 J/AJ/116/2728 J/AJ/116/2738 J/MNRAS/332/422

Scott D.: J/MNRAS/358/149 J/MNRAS/372/1621 J/MNRAS/376/1073 Scott D.K.: I/100A I/100A SCOTT D.K.: I/39 Scott F.P.: I/100A Scott J.: J/ApJ/571/665 J/ApJS/130/37 J/ApJS/140/143 Scott M.P.: J/A+A/446/1185 J/A+A/475/765 Scott P.F.: J/MNRAS/341/1066 J/MNRAS/342/915 J/MNRAS/360/340 SCOTT P.F.: VII/53 VII/53 VIII/4 VIII/4 Scott S.E.: J/MNRAS/370/1057 J/MNRAS/372/1621 Scott W.K.: J/ApJ/616/110 J/ApJS/131/95 J/ApJS/155/33 J/ApJS/90/173 J/PASJ/52/997 VIII/40 Scoville N.: J/ApJ/599/1049 Scoville N.Z.: J/AJ/128/1974 J/ApJ/315/104 J/ApJ/633/871 J/ApJ/644/100 J/ApJS/63/821 VII/114 Scowen P. A.: J/PASP/108/682 Scowen P.: J/AJ/113/1001 Scowen P.A.: J/AJ/104/92 J/AJ/117/1700 Scranton R.: J/AJ/127/1811 Scuderi S.: J/A+A/377/123 J/A+A/420/683 J/AJ/121/2020 J/AJ/121/327 Scuflaire R.: J/A+A/415/241 J/A+A/463/243 Seager S.: J/A+A/275/484 J/ApJ/582/1123 Seal J.: J/A+AS/147/169 J/ApJS/98/369 SEATON M.J.: VI/80 Seaton M.J.: VI/89A Sebata K.: J/PASJ/52/631 Sebo K.M.: J/AJ/108/932 J/ApJ/512/48 J/ApJ/523/540 J/ApJ/529/723 J/ApJS/142/71 Sebo P.R.: J/ApJ/449/164 Seckel D.: J/other/APh/26.282 Sedano J.L.: J/A+A/340/196 SEDANO J.L.: J/A+AS/93/189 Sedgwick N.: J/MNRAS/351/1290 Sedlmayr E.: J/A+A/361/641

Seeberger R.: J/A+A/286/17 J/A+AS/110/269 J/A+AS/117/1 J/A+AS/126/247 J/A+AS/127/101

J/AJ/107/2101 Seetha S.: J/A+A/455/303 J/ApJS/156/217 Segall M.: J/MNRAS/375/831 SEGATO C.: II/130 SEGEEWISS W.: J/A+A/301/396 Seggewiss W.: J/A+A/331/949 J/A+A/376/745 J/A+A/381/884 SEGGEWISS W.: J/A+AS/112/367 Seggewiss W.: J/A+AS/139/425 Segransan D.: J/A+A/364/665 J/ApJ/620/961 Seguel J.: J/MNRAS/358/1215 J/MNRAS/377/300 Seibert M.: J/AJ/130/825 J/ApJS/164/38 Seidelmann P.K.: J/AJ/112/2319 Seifert W.: J/A+A/398/49 Seigar M.: J/MNRAS/372/1621 Seigar M.S.: J/A+A/414/23 J/MNRAS/373/781 Seigmund W.: J/AJ/121/2331 Seiradakis J.H.: VII/8 Seitz S.: J/A+A/398/49 J/A+A/418/885 J/A+A/420/97 J/A+A/421/41 J/A+A/445/423 J/A+A/465/375

J/MNRAS/372/1425 Seitzer P. Irwin M.J.: J/ApJS/90/83 Seitzer P.: J/A+A/396/473 J/A+A/398/479 J/A+A/445/503 SEITZER P.: J/AJ/106/986 Seitzer P.: J/AJ/110/652 J/AJ/112/2408 J/AJ/130/1111 Seitzer P. O.: J/AJ/115/708 Seitzer Patrick: J/ApJ/347/251 Sekguchi K.: J/MNRAS/362/1371 Sekido M.: J/PASJ/52/631 Sekiguchi K.: J/ApJ/634/861 J/MNRAS/274/1194 J/MNRAS/316/326 J/MNRAS/321/77 J/MNRAS/342/802

J/MNRAS/372/1621 J/MNRAS/372/741 J/PASJ/55/1079 J/PASJ/57/165 Sekiguchi M.: J/AJ/123/1807 J/ApJS/137/279 J/MNRAS/330/137 J/MNRAS/331/609 J/MNRAS/357/1357

J/PASJ/55/1079 J/PASJ/55/771 Sekiguchi T.: J/A+A/380/347 SELAM S.: J/A+AS/106/373 Selam S.: J/MNRAS/371/703 Selam S.O.: J/MNRAS/374/664 J/other/PASA/23.154 Seljak U.: J/ApJS/163/80 Sellgren K.: J/ApJ/470/864 J/ApJ/597/323 J/ApJS/143/73 Sellier A.: J/A+AS/137/451 Selman F.: J/A+A/341/98 Selman F.J.: J/A+A/443/851 Selvelli P.: J/A+AS/124/425 Selvelli P.L.: J/A+A/381/507 Sembach K.R.: J/ApJ/480/568 J/ApJS/136/631 J/ApJS/146/1 J/ApJS/146/125 J/ApJS/155/351 Semenova T.A.: J/AZh/84/227 Semkov E.H.: VI/90 Sen A.K.: J/A+AS/141/175 Senay M.C.: J/AJ/109/1269 Senda A.: J/ApJ/635/214 Sen’e D.S.: J/AZh/71/762 Seo S.H.: J/other/APh/26.282 Serabyn E.: J/ApJ/581/258 Serafino A.: J/A+AS/118/1 Sergeev A.: J/A+A/383/724 Sergeev A.V.: J/other/KFNT/15.483 Sergeev E.A.: J/A+A/371/79 Sergeev S. G.: J/ApJ/510/659 J/AZh/71/200 Sergeev S.G.: J/A+A/320/405 J/A+A/324/904 J/A+A/369/57 J/A+A/371/79 J/A+A/437/87 J/ApJ/622/129

J/ApJS/115/185 J/AZh/71/189 J/other/Ap/48.304 J/PAZh/23/409 J/PAZh/27/942 Sergeeva E. A.: J/ApJ/510/659 Sergeeva E.A.: J/A+A/437/87 J/ApJ/622/129 J/ApJS/115/185 J/other/Ap/48.304 Sergeeva T.: I/261

Serjeant S.: J/AJ/127/3075 J/AJ/129/1198 J/MNRAS/325/1173 J/MNRAS/343/293 J/MNRAS/344/887 J/MNRAS/351/1290 J/MNRAS/355/97 J/MNRAS/358/333 J/MNRAS/358/397 J/MNRAS/370/1057 J/MNRAS/372/1621 J/MNRAS/375/115

Serkowitsch E.: J/A+AS/140/29 Serlemitsos P.J.: J/ApJ/253/485 J/ApJS/96/303 Serote Roos M.: J/A+A/429/101 J/A+AS/117/93 Serra P.: J/A+A/434/53 J/A+A/453/493 Serrano A.: J/ApJS/117/319 Sestito P.: J/A+A/465/185 Seth A.: J/AJ/112/2769 Seth A.C.: J/ApJ/581/325 Setia Gunawan D.Y.A.: J/A+A/331/949 J/ApJS/149/123 Seunarine S.: J/other/APh/26.282 Sevastyanenko T.: J/AJ/130/2804 SEVENSTER M.: II/200 Sevenster M.: J/A+A/390/501 J/A+A/390/511 Sevenster M.N.: J/A+A/366/481 J/A+AS/122/79 J/A+AS/124/509 J/AJ/123/2772 Severgini P.: J/A+AS/137/495 Severgnini P.: J/A+A/409/79 J/A+A/428/383 J/A+A/476/1191 Sevilla I.: J/A+A/390/39 J/A+A/410/813 Sevre F.: J/A+AS/122/489 J/A+AS/125/399 Seward F.D.: IX/27 J/ApJ/589/509 J/ApJS/73/781 VII/19 Sewilo M.: J/ApJS/154/553 Seyfert C.K.: II/221A Seymour D.M.: J/AJ/123/1023 Seymour N.: J/A+A/475/791 J/MNRAS/352/131 J/MNRAS/362/1371 J/MNRAS/369/156 J/MNRAS/378/995 Sezer C.: J/A+A/409/959 Sfeir D.: J/A+A/346/785 J/A+A/411/447 Sguera V.: J/ApJ/636/765 J/ApJS/170/175 Shafer R.A.: J/ApJ/253/485 SHAFFER D.B.: I/151 Shaffer D.B.: J/AJ/108/1163 J/AJ/110/880 Shaimukhametov R.: J/A+A/375/670 Shakeshaft J.R.: VIII/1A Shakhbazian R.K.: VII/89B VII/89B VII/89B Shakhovskaya N.I.: II/55 J/A+AS/139/555 Shakhovskoj D.: J/A+A/448/1075 Shakhovskoy D.N.: J/A+A/327/145 J/AZh/78/60 Shakhovskoy: J/AZh/77/420 Shakhovskoy N.M.: J/A+A/454/277 Shakht N.A.: I/292 Shan H.G.: J/other/NewA/11.557 Shane C.D.: I/11 Shane N.S.: J/A+A/414/23 Shane W.W.: J/A+AS/61/451 Shang Z.: J/AJ/117/2757 J/MNRAS/375/989 Shanks T.: II/268 J/ApJ/505/L7 J/MNRAS/235/827 J/MNRAS/236/207 J/MNRAS/243/1 J/MNRAS/251/482

J/MNRAS/276/33 J/MNRAS/300/417 J/MNRAS/306/592 J/MNRAS/328/805 J/MNRAS/335/673 J/MNRAS/349/135 J/MNRAS/352/91 J/MNRAS/372/425 J/MNRAS/375/68 VII/223 VII/241

Shao M.: J/AJ/130/2866 J/ApJ/636/1020 Shao Zh.-Y.: J/A+AS/131/89 Shaparenko E.F.: J/other/KFNT/12.20 Shapley A.E.: J/ApJ/592/728 Shapovalov A.N.: III/230 Shapovalova A. I.: J/ApJ/510/659 Shapovalova A.I.: J/A+A/371/79 J/A+A/376/775 J/A+A/422/925 J/ApJS/115/185 SHAPOVALOVA A.I.: VII/172 Shara M.: J/AJ/112/2408 J/AJ/117/206 J/AJ/127/2771 J/MNRAS/357/645 J/PASP/118/1238

Shara M.M.: I/298 J/AJ/106/1639 J/AJ/120/2101 J/AJ/124/1190 J/AJ/125/1598 J/ApJ/625/156

J/PASP/109/998 V/110 V/123A V/79 V/94 Sharaf M.A.: J/other/NewA/8.645 Share G. H.: J/ApJS/120/409 Sharif L.: J/A+AS/135/503

Sharina M.E.: J/A+A/396/473 J/A+A/398/479 J/A+A/442/85 J/A+AS/119/499 J/MNRAS/372/1259

VII/219 Sharma A.K.: J/A+A/376/333 J/A+A/376/356 J/A+AS/117/557 Sharma M.: J/other/BASI/30.911 Sharma S.: J/MNRAS/379/1237 J/other/BASI/35.53 Sharon K.: J/ApJ/621/53 Sharov A.S.: J/PAZh/23/113 VI/123 Sharp R.: J/A+A/476/1191 J/MNRAS/372/425 Sharpee B.: J/AJ/123/3216 J/ApJS/149/157 Sharples R.: J/MNRAS/246/54 J/MNRAS/355/608 Sharples R.M.: J/A+A/448/155 J/AJ/118/752 J/ApJ/497/188 J/ApJS/110/213 J/MNRAS/341/1093 Sharpless S.: VII/20 Shastri P.: J/A+A/425/99 Shatskii N.I.: J/PAZh/24/918 Shatsky N.: J/A+A/380/238 J/A+A/382/92 J/A+AS/139/69 Shatsky N.I.: J/A+A/409/245 J/PAZh/24/307 Shaver P.: J/A+A/425/367 J/A+AS/119/265 Shaver P.A.: J/A+A/376/1123 J/A+A/386/97 J/A+A/399/469 J/other/AuJPA/14.1 J/other/AuJPA/14.133

J/other/AuJPA/14.77 Shavrina A.: J/A+A/430/255 SHAW G.: J/MNRAS/253/686 Shaw R.A.: J/ApJ/622/294 J/PASP/109/289 Shaw S.E.: J/ApJ/607/L33 J/ApJ/636/765 Shaya E.J.: J/ApJ/448/179 J/ApJ/459/L27 J/ApJS/80/479 Shchegolev D.E.: III/201 Shchegoleva N.V.: J/AZh/77/499 Shcherbakov A.G.: J/A+A/309/655 Shcherbakov V.A.: J/A+A/341/553 SHCHERBAKOV V.A.: J/A+AS/107/9 Shcherbakova Z.A.: J/A+A/309/655 Shcherbanovskij A.L.: VII/77 VII/82A Shectman S.: J/AJ/132/137 Shectman S.A.: J/AJ/103/1987 J/AJ/103/267 J/AJ/108/538 J/AJ/125/2502 J/AJ/132/85 J/AJ/88/1285

J/AJ/90/2089 J/AJ/95/284 J/ApJ/314/493 J/ApJS/130/237 J/ApJS/76/1001 VII/203 SHEERAN M.: J/PASP/106/486 Sheets H.A.: J/ApJ/640/1018 Sheffer E.K.: J/A+AS/114/341 J/AZh/79/1087 Shefler T.: J/AJ/117/981 Sheikina T.A.: J/A+AS/147/5 Sheldon E.: J/AJ/119/2598 Shelton I.: J/AJ/113/1711 Shelton R.G.: J/AJ/115/2587 Shemi A.: J/MNRAS/287/293 Sheminova V. A.: J/A+A/329/721 Shemmer O.: J/MNRAS/324/580 Shen J.: J/AJ/131/84 Shen K.X.: J/A+A/383/296 J/A+A/391/775 J/A+A/422/377 J/A+A/454/379 J/MNRAS/376/1707 Shen Z.-Q.: J/ApJ/616/110 J/ApJ/640/982 J/ApJS/155/33 J/PASJ/52/997

Shenavrin V.I.: III/209 J/A+A/394/617 J/A+AS/92/1 J/AZh/71/618 J/AZh/79/1118 J/PAZh/25/860

J/PAZh/26/695 J/PAZh/27/393 J/PAZh/30/605 J/PAZh/32/866 Sheng M.: J/A+A/466/255 Shengquan L.: J/A+AS/114/337 Shepherd C.W.: J/ApJ/565/208 J/ApJS/129/475 Shepherd D.S.: J/ApJ/602/850 Shepherd M.C.: VIII/72 Sheppard D.C.: J/MNRAS/328/17 Sheppard S.S.: J/AJ/127/3023 Sherry W.: J/AJ/112/2408 Sherry W.H.: J/AJ/128/2316 Sheth K.: J/ApJ/633/871 Sheth R.K.: J/AJ/125/1817 J/AJ/130/968 J/AJ/131/1288 J/AJ/131/2018 Shetrone M.D.: J/A+A/459/423 J/AJ/113/279 J/AJ/126/2954 J/AJ/128/1177 Shevchenko V. S.: J/AZh/71/572 SHEVCHENKO V.S.: J/A+A/282/L9 Shevchenko V.S.: J/A+A/350/513 J/A+A/374/1049 J/A+A/461/183 SHEVCHENKO V.S.: J/A+AS/103/183 J/A+AS/107/9 Shevchenko V.S.: J/A+AS/124/33 J/A+AS/139/513 J/AJ/118/1043 J/MNRAS/310/210 SHEVELEV Yu.G.: V/56 V/57 Shevelev Yu.G.: V/58 SHEVELEV Yu.G.: V/89 Shi H.M.: J/A+AS/100/243 Shi J.R.: J/A+A/423/683 J/A+A/425/671 Shi J.-R.: J/PASJ/58/389 Shi Y.: J/A+A/426/503 Shibai H.: J/A+A/365/409 J/A+A/423/579 Shibata K.M.: J/PASJ/52/631 J/PASJ/52/997 Shields G.A.: J/ApJ/513/168 Shields J.: J/ApJS/161/9 Shields J.C.: J/AJ/127/105 J/ApJ/641/140 Shih D.: J/ApJS/163/80 Shimasaku K.: J/ApJ/634/861 J/ApJ/648/7 J/ApJS/137/279 J/MNRAS/357/1357 J/PASJ/55/1079

J/PASJ/57/165 J/PASJ/58/313 Shimizu R.: J/PASJ/52/631 Shimoikura T.: J/PASJ/52/631 Shin M.: J/A+A/437/61 J/A+A/445/69 Shine S.A.: J/A+A/423/235 Shioya Y.: J/ApJ/648/7 J/PASJ/57/165 J/PASJ/57/881 J/PASJ/58/313 Shipman R.F.: J/AJ/126/1423 Shirasaki Y.: J/other/GCN/1070.1 Shiroya R.: J/ApJS/112/245 Shishkina V.N.: I/78 Shkhagosheva Z.U.: J/A+A/385/87 J/A+A/422/627 J/A+AS/105/503 J/A+AS/140/287 Shkodrov V.: J/A+AS/125/399 Shkutov V.D.: I/242 Shobbrook R.R.: J/A+A/374/1056 Shokin Y.A.: J/PAZh/22/847 Shokin Yu.A.: J/PAZh/20/693 J/PAZh/21/509 J/PAZh/22/269 J/PAZh/24/93 Sholukhova O.: J/A+AS/140/309 Sholukhova O.N.: J/PAZh/25/18 Shopbell P.: J/ApJ/471/L5 J/ApJ/538/29 SHOR V.A.: I/145 I/183B I/184C Shor V.A.: I/204 SHOR V.A.: I/217 I/233 Shor V.A.: I/245 Shore S.N.: III/89 Shott P.: J/PASP/99/739 Shower C.A.: J/A+A/424/993 Shpil’kina D.A.: J/A+A/320/79 Shporer A.: J/MNRAS/370/1429 Shrader C.: J/ApJS/99/223 Shrader C.R.: J/ApJS/154/585 Shriver S.K.: J/ApJS/101/259 Shubert R.: J/ApJS/127/79 Shugarov S.: J/A+A/323/113 J/A+A/366/100 Shugarov S.Y.: J/PAZh/27/187 Shugarov S.Yu.: II/139B II/140 J/AZh/80/535 J/PAZh/28/764 J/PAZh/29/858 Shukla H.: J/ApJS/120/147 Shull J.M.: J/ApJS/146/1 J/ApJS/146/125 Shupe D.: J/A+A/467/565 J/AJ/129/1198 J/AJ/129/2074 J/MNRAS/371/1891 Shupe D.L.: II/255 J/A+A/467/73 J/AJ/125/2411 J/AJ/131/250 J/PASP/113/10 Shuping R.Y.: J/AJ/130/1763 Siana B.: J/A+A/467/565 J/AJ/129/2074 Sibille F.: J/A+A/350/883 J/A+A/357/219 J/A+A/372/173 J/A+A/420/945 J/A+A/421/623 Sicilia-Aguilar A.: J/AJ/128/805 J/AJ/129/363 J/AJ/129/856 J/AJ/130/188 J/ApJ/638/897 Siciliano L.: J/AJ/112/2408 Sidoli L.: J/A+A/368/835 Sidorenkov V.N.: J/A+AS/114/341 Siebenmorgen R.: J/A+A/365/545 J/A+A/473/369 Siebert A.: J/A+A/398/141 SIEBERT J.: J/A+A/281/355 J/A+A/285/812

Siebert J.: J/A+A/323/739 J/A+A/330/108 J/A+A/356/445 J/A+A/363/141 J/A+AS/106/303

J/A+AS/109/147 J/A+AS/122/235 J/ApJ/525/127 Siebert K.W.: J/MNRAS/284/265 Siegel M.H.: J/AJ/120/284 J/AJ/121/935 J/AJ/125/1352 J/ApJ/623/181 Siegler N.: J/AJ/131/1574 Siegman B.: VII/75 VIII/56 Siegman B.C.: J/MNRAS/235/1313 Siegmund O.H.W.: J/AJ/130/825 J/ApJ/619/L95 J/ApJS/164/38 Siegmund W.A.: J/AJ/121/2308 J/AJ/123/567 Siems M.: J/A+A/390/39 J/A+A/406/L9 J/A+A/410/813 Sievers A.: J/A+AS/102/611 J/A+AS/122/271 Sijbring D.: J/A+AS/116/15 Sikkema G.: J/A+A/392/741 Silbermann M.A.: J/ApJ/514/614 Silbermann N.: J/AJ/123/3216 SILBERMANN N.A.: J/AJ/107/1764 J/AJ/107/679 Silbermann N.A.: J/AJ/110/704 J/ApJ/507/655 J/ApJ/508/491 J/ApJ/512/48 J/ApJ/521/155 J/ApJ/523/540

J/ApJ/529/723 Silchenko O.K.: J/AZh/71/706 Silk J.: J/A+A/432/381 J/ApJ/475/445 Sillanpaa A.: J/A+AS/129/577 J/A+AS/130/109 J/A+AS/143/357 Silva B.: J/A+A/468/353 Silva D.R.: III/166 J/A+A/406/75 J/A+A/438/521 J/AJ/126/2644 Silva L.: J/A+A/364/517 J/MNRAS/351/1290 J/MNRAS/372/1621 Silva R.P.: J/A+A/458/895 Silverman J.: J/ApJ/645/955 Silverman J.D.: J/ApJ/618/123 J/ApJ/644/829 J/ApJS/150/19 J/ApJS/161/21 Silverstone M.: J/PASP/118/1690 Silverstone M.D.: J/ApJ/555/932 Silvester J.: J/A+A/475/1053 Silvestri A.: J/other/APh/26.282 Silvestri N.M.: J/AJ/124/1118 J/AJ/126/1023 J/AJ/128/426 J/AJ/129/2428 J/AJ/131/1674 J/ApJ/607/426 Silvotti R.: J/A+A/428/339 J/A+A/459/557 Sim S.A.: J/MNRAS/361/1102 Simard L.: J/A+A/427/397 J/A+A/444/365 J/AJ/130/2043 J/ApJ/507/585 J/ApJ/549/172 J/ApJ/571/136

J/ApJ/599/1006 J/ApJ/630/206 J/ApJS/142/1 J/ApJS/157/175 J/ApJS/159/41 J/MNRAS/361/109 J/MNRAS/371/1777 J/PASP/109/883

Simard-Normandin M.: J/ApJS/45/97 Simic D.: J/ApJS/129/435 Simic Z.: J/A+A/417/375 J/A+A/425/1147 J/A+A/441/391

Simien F.: J/A+A/309/749 J/A+A/366/68 J/A+AS/117/417 J/A+AS/118/111 J/A+AS/122/521 J/A+AS/126/15 J/A+AS/126/519 J/A+AS/127/117 J/A+AS/131/287 J/A+AS/133/317 J/A+AS/136/519 J/A+AS/145/263

Simien: J/A+A/384/371 J/A+AS/136/509 J/A+AS/139/483 J/A+AS/31/439 Simmerer J.: J/ApJ/617/1091 J/ApJ/645/613 Simon A.J.B.: J/MNRAS/238/1171

Simon G.: II/243 J/A+A/349/L69 J/A+A/362/215 J/A+A/363/62 J/A+A/380/590 J/A+A/400/903

J/A+A/401/959 J/A+A/403/141 J/A+A/446/551 J/ApJ/552/289 Simon J.L.: J/A+A/329/329 SIMON J.L.: VI/66 Simon M.: I/123 SIMON M.: I/142 Simon M.: J/AJ/117/1375 SIMON N.R.: J/ApJ/412/183 Simon P.: J/A+A/441/905 Simon R.: J/ApJ/639/227 Simon R.S.: J/AJ/114/2822 Simon T.: J/A+A/375/989 J/AJ/129/829 J/AN/319/215 SIMON T.: J/ApJ/408/305 Simon T.: J/PASP/109/759 Simon V.: J/A+A/309/521 J/A+A/366/100 Simon-diaz S.: J/A+A/437/467 Simonia I.A.: J/PAZh/30/946 Simons L.W.J.: VIII/20 Simonsen M.: J/other/IBVS/5526 Simpson C.: J/ApJ/634/861 J/MNRAS/372/1621 J/MNRAS/372/741 J/MNRAS/373/1531 J/MNRAS/379/1343 Simpson J.P.: J/ApJS/162/346 Sinachopoulos D.: J/A+A/402/1157 J/A+A/472/1055 SINACHOPOULOS D.: J/A+AS/100/517 Sinachopoulos D.: J/A+AS/112/291 J/A+AS/119/483 J/A+AS/124/353 J/A+AS/136/525 J/A+AS/139/69

J/MNRAS/259/563 Sinamyan P.: J/other/Ap/46.475 Sinamyan P.K.: J/other/Ap/46.46 Sinelnikov I.Ye.: J/A+A/420/447 Singh H.P.: J/ApJS/152/201 J/ApJS/152/251 Singh P.D.: J/A+AS/120/301 Sinha S.: J/ApJS/156/217 Sinvhal S.D.: II/40 Siodmiak N.: J/A+A/373/1032 J/A+A/405/627 J/A+A/469/799 Sion E.M.: III/129 III/210 III/235A J/ApJ/571/512 J/ApJS/119/207 Siopis C.: J/AJ/129/2138 Sipahi E.: J/MNRAS/376/573 Sirianni M.: J/AJ/128/1990 J/ApJ/612/202 J/ApJ/621/53 J/ApJS/150/1 Sirk M.M.: J/AJ/107/751 J/AJ/108/1843 Sirko E.: J/AJ/127/899 Sivagnanam P.: J/A+A/267/515 J/A+A/277/453 J/A+A/394/975 Sivakoff G.R.: J/ApJ/599/218 J/ApJ/617/262 Sivan J.P.: J/A+A/375/L27 J/A+A/410/1039 J/A+A/410/1051 J/A+A/414/351 SIVAN J.-P.: J/A+AS/107/385 Siviero A.: J/A+A/434/1107 J/A+A/452/567 S.J.: J/MNRAS/325/1002 Sjouwerman L.O.: J/A+A/391/967 J/A+A/393/115 J/A+A/418/103 J/A+A/435/575 J/A+AS/128/35 Skelton B.P.: J/PASP/111/685 Skelton R.T.: J/ApJS/127/79 Skiff B. A.: III/233B Skiff B.: II/221A II/277 IV/23 J/other/IBVS/4431

Skiff B.A.: B/mk IV/29 J/A+A/370/496 J/other/IBVS/4433 J/other/IBVS/4441 J/other/IBVS/4459

J/other/IBVS/4678 J/other/IBVS/5185 Skiff B.A: J/other/IBVS/5187

Skillen I.: J/A+A/330/515 J/A+A/467/785 J/ApJS/156/227 J/MNRAS/362/1006 J/MNRAS/372/1117

J/MNRAS/379/647 J/MNRAS/379/816 Skillman D.R.: J/ApJS/115/185 Skillman E.D.: J/AJ/123/3154 J/AJ/125/1261 J/ApJ/513/168 Skinner C. J.: J/ApJS/122/221 Skinner G.K.: J/ApJ/633/L77 J/PAZh/28/323 Skinner S.: J/ApJ/598/375 Skinner S.L.: J/A+A/468/353 J/ApJ/613/393 Skoda P.: J/A+A/309/521 J/A+A/416/669 J/A+A/419/607 J/A+A/446/583 J/A+A/455/1037 J/A+A/459/849

J/A+A/463/1061 J/A+A/463/233 Skopal A.: J/A+A/318/53 J/A+A/348/184 J/A+A/367/199 Skrutskie M.: J/AJ/125/2521

Skrutskie M.F.: II/246 J/AJ/116/1816 J/AJ/121/3160 J/AJ/124/1001 J/AJ/124/1570 J/AJ/128/245

J/ApJ/476/311 J/ApJ/564/421 Skuljan J.: J/MNRAS/330/137 Slane P.: J/ApJ/615/887 Slater M.: V/62 V/77 Slavin S. D.: J/AJ/115/708 Slawson R. W.: J/AJ/117/286 Slawson R.W.: J/ApJS/82/117

Slechta M.: J/A+A/367/236 J/A+A/416/669 J/A+A/419/607 J/A+A/446/583 J/A+A/455/1037

J/A+A/459/849 J/A+A/463/1061 J/A+A/463/233 J/A+A/464/263 Slee O.B.: J/MNRAS/235/1313 VIII/35 VIII/50 VIII/50 Slesnick C.L.: J/ApJ/576/880 SLETTEBAK A.: III/14 Slettebak A.: III/15 III/76A

Slezak E.: J/A+A/430/19 J/A+A/461/81 J/A+AS/111/265 J/A+AS/119/329 J/A+AS/120/357 J/A+AS/122/409 J/A+AS/128/67 J/A+AS/129/281 J/A+AS/139/525 J/A+AS/139/559 J/AJ/110/463 J/other/PASA/16.113

SLEZAK E.: VII/159

Slijkhuis R.: J/A+A/379/436 J/A+A/379/740 J/A+A/384/81 J/A+A/392/741 J/A+A/452/119

J/A+A/456/881 J/A+A/462/553 J/A+AS/137/75 J/A+AS/137/83 Slinglend K.: J/ApJS/115/1 Slinglend K.A.: J/ApJ/520/491 Slivan S.: J/A+A/465/331 Sloan G.C.: J/ApJS/119/141 J/MNRAS/376/313 Slosar A.: J/MNRAS/341/1066 J/MNRAS/360/340 Sluis A.P.N.: J/AJ/131/2089 Sluse D.: J/A+A/433/757 J/A+A/441/915 Slysh V.: J/PASJ/52/997 Slysh V.I.: J/A+AS/124/85 J/A+AS/134/115 J/AZh/71/37 J/AZh/72/22 Slyz A.: J/ApJ/412/479 S.M.: J/A+A/372/879

Smail I.: J/ApJ/497/188 J/ApJ/549/820 J/ApJ/581/844 J/ApJ/591/53 J/ApJ/614/671 J/ApJ/616/71 J/ApJ/622/772 J/ApJ/634/977 J/ApJS/110/213 J/ApJS/122/51 J/MNRAS/323/839 J/MNRAS/327/588 J/MNRAS/330/1 J/MNRAS/335/10 J/MNRAS/337/256 J/MNRAS/346/1125 J/MNRAS/358/233 J/MNRAS/365/439 J/MNRAS/366/645 J/MNRAS/372/1621 J/MNRAS/379/1343

Smale A.P.: J/ApJS/96/303 Small T.: J/AJ/130/825 J/ApJ/619/L95 J/ApJS/164/38 Small T.A.: J/PASP/106/646 Smalley B.: J/A+A/271/515 J/A+A/328/349 J/MNRAS/363/529 J/MNRAS/364/272 Smart R.L.: J/A+A/448/579 J/A+AS/121/119 Smartt S.J.: J/A+A/334/987 J/A+A/367/86 J/A+A/373/608 J/A+A/437/467 J/A+A/456/623 J/A+A/457/265

J/A+A/466/277 J/A+A/471/625 J/A+AS/139/231 J/MNRAS/351/1071 SMATHERS H.W.: VII/80

Smecker-Hane T.A.: J/AJ/120/1808 J/AJ/129/1465 J/ApJ/471/694 J/ApJS/102/289 J/ApJS/113/1

J/MNRAS/378/318 J/PASP/110/533 Smee S.: J/AJ/121/2308 J/AJ/121/2331 J/AJ/123/567 Smekhov M.G.: J/A+A/382/118 Smetanka J.J.: J/ApJS/109/45 Smette A.: J/A+A/473/369 SMIRNOV O.: J/A+AS/112/439 Smirnov O.: J/A+AS/119/191 J/A+AS/125/1 J/AJ/112/2004 J/AJ/112/2020 J/AJ/114/189 J/AJ/116/2415 Smirnov O.M.: J/PAZh/21/596 Smirnov Yu.M.: J/AZh/79/932 J/AZh/80/266 J/AZh/81/33 Smirnova Yu.V.: J/AZh/78/922 Smith A.J.: J/other/APh/26.282 Smith A.M.: J/AJ/108/1350 J/AJ/112/601 J/AJ/113/1011 J/AJ/116/180 J/AJ/116/941 SMITH A.M.: J/ApJ/413/604 J/ApJ/425/122 Smith A.M.: J/ApJ/438/181 J/ApJ/446/622 J/ApJS/104/287 J/ApJS/98/595 SMITH A.M.: VII/48 Smith B.J.: J/AJ/123/1411 J/AJ/131/612 J/ApJS/154/673 Smith C.: J/AJ/116/1 V/31A Smith C.A.: I/138 SMITH C.A.: I/160 Smith C.A.: I/161 Smith D.A.: J/AJ/122/3046 J/AJ/127/2436 J/ApJ/591/138 J/ApJ/603/503 Smith D.M.: J/A+A/407/1039 Smith E.: J/A+A/437/789 J/AJ/127/2738 J/AJ/130/1693 Smith E.O.: J/AJ/111/1596 J/AJ/114/1471 Smith E.P.: J/AJ/108/1350 J/AJ/112/601 SMITH E.P.: J/ApJ/413/604 Smith E.P.: J/ApJ/438/181 J/ApJS/104/287 J/ApJS/98/595 Smith G.H.: J/AJ/113/279 J/AJ/125/197 J/AJ/127/1588 J/AJ/97/1360 Smith G.P.: J/ApJ/627/32 J/ApJ/634/977 J/MNRAS/323/839 J/MNRAS/330/1 J/MNRAS/379/1343 Smith H.: J/AJ/123/3216 Smith H.A.: J/A+A/445/971 SMITH H.A.: J/AJ/107/1764 J/AJ/107/679 Smith H.A.: J/ApJ/640/L43 J/PASP/117/1325 Smith H.E.: J/A+A/467/565 J/A+A/467/73 J/AJ/129/1198 J/AJ/129/2074 SMITH H.E.: VII/17 Smith Horace A.: J/AJ/110/704 Smith J.: J/PASJ/52/997 Smith J.A.: J/AJ/119/2598 J/AJ/126/1023 J/AJ/127/1883 J/AJ/127/2210 J/AJ/129/2062 J/AJ/130/873

J/AJ/131/1674 J/ApJ/607/426 J/MNRAS/372/425 Smith J.D.: J/A+A/469/575 J/ApJ/633/871 Smith J.D.T.: J/AJ/132/231 Smith K.C.: J/A+A/274/335 Smith M.A.: J/A+A/458/569 Smith M.C.: J/MNRAS/370/435 Smith M.D.: J/A+A/426/171 J/A+A/447/609 J/A+A/470/605 Smith M.G.: J/ApJ/618/123 J/ApJ/645/955 Smith N.: J/A+A/419/215 J/AJ/130/1763 Smith P.L.: J/A+A/377/361 Smith P.S.: J/AJ/130/1418 J/AJ/132/676 J/ApJ/569/23 J/ApJ/618/123 Smith R.: J/MNRAS/352/478 J/MNRAS/357/819 Smith R.A.: J/A+A/383/580 Smith R.C.: J/AJ/112/2408 J/AJ/130/2453 J/AJ/131/1648 J/AJ/131/2478 J/ApJ/594/1 J/ApJ/602/571

J/ApJ/618/123 J/ApJ/645/955 J/ApJS/165/307 Smith R.J.: II/268 J/AJ/128/1558 J/ApJ/505/L7 J/MNRAS/306/592 J/MNRAS/313/469 J/MNRAS/327/249

J/MNRAS/327/265 J/MNRAS/328/805 J/MNRAS/335/673 VII/223 VII/241 Smith R.K.: J/ApJ/459/686 Smith R.M.: J/A+A/476/59 J/MNRAS/304/297 VII/191 Smith S.B.: VII/171 Smith T.E.: J/AJ/119/486 J/AJ/120/2747 Smith T.L.: J/ApJS/143/377 Smith T.R. Mateo M.: J/ApJ/521/577 Smith V.V.: J/ApJS/151/313 Smith W.B.: IV/22 SMITH-MOORE M.: III/133 Smoker J.V.: J/A+A/419/1123 J/MNRAS/367/1478 Smoot G.: J/A+A/397/213 J/ApJ/500/554 SMRIGLIO F.: II/130 Smriglio F.: J/A+A/388/149 J/A+A/458/767 J/BaltA/12/323 J/BaltA/7/369 Snedden S.A.: J/A+A/444/643 J/AJ/123/3409 J/AJ/124/1651 J/AJ/131/1674 J/AJ/132/676 J/ApJS/115/185

J/PASJ/55/771 SNEDEN C.: J/A+A/287/927

Sneden C.: J/A+A/354/169 J/A+A/356/238 J/AJ/112/545 J/AJ/113/279 J/AJ/130/2804 J/AJ/132/85

J/ApJ/530/783 J/ApJ/562/528 J/ApJ/617/1091 J/ApJ/645/613 J/ApJS/162/227 J/PASP/111/1233 Snegireva E. A.: J/A+A/446/785 Snell R.L.: J/ApJ/551/852 Snellen I.A.G.: J/A+A/402/171 J/A+A/435/863 J/MNRAS/329/700 Snider S.: J/ApJ/562/528 Snigula J.: J/A+A/398/49 J/A+A/421/41 Snir Y.: J/AJ/121/2308 J/AJ/121/2331 J/AJ/123/567 Snodgrass C.: J/A+A/444/287 Snodgrass H.B.: J/AJ/106/2058 Snoek L.C.: J/A+AS/116/257 Snow M.: J/AJ/116/941 Snow T.P.: II/149 III/77 J/ApJS/95/163 Snowden S.L.: J/A+A/316/147 J/ApJ/561/189 Snyder J.A.: J/AJ/127/1158 SNYDER L.E.: VIII/12 Soares D.S.L.: J/A+AS/110/371 Soares J.: J/A+A/404/223 Soares: J/A+A/400/533 Sobczak G.J.: J/AJ/131/527 Sobczynska D.: J/A+A/410/813 Sobolev A.M.: J/A+A/434/201 Soboleva N.S.: J/A+AS/101/447 J/A+AS/87/1 J/A+AS/91/337 J/AZh/71/684 J/AZh/72/613 J/AZh/73/835

J/AZh/79/589 J/other/BSAO/40.5 VIII/33 Sobrito G.: J/A+AS/130/109 J/A+AS/132/361 J/A+AS/144/481 Sobzynska D.: J/A+A/406/L9 Socas-Navarro H.: J/A+A/465/339 Sochting I.K.: J/MNRAS/369/1334 Soderberg A.M.: J/AJ/131/527 J/PASP/111/1233 Soderblom D.: J/PASP/118/1690 SODERBLOM D.R.: J/AJ/105/226 J/AJ/105/2299 Soderblom D.R.: J/AJ/106/1059 SODERBLOM D.R.: J/AJ/106/1080 Soderblom D.R.: J/AJ/111/439 SODERBLOM D.R.: J/ApJS/85/315 Sodor A.: J/A+A/430/1049 J/A+A/445/617 J/A+A/469/1033 Sodre Jr. L.: J/A+A/407/31 Sodre L.: J/A+A/258/243 SODRE L.: J/MNRAS/259/233 Sodre L. Jr.: J/A+AS/114/565 Sodre L.Jr.: J/A+A/377/428 Sodre L.Jr: J/MNRAS/274/1107 Soederhjelm S.: J/A+A/341/121 Soegiartini E.: J/A+AS/107/235 J/A+AS/114/487 J/A+AS/134/87 Soffitta P.: J/A+A/455/813 J/A+A/473/347 Soffner T.: J/A+A/306/9 Sofia S.: J/AJ/127/1158 Sofue Y.: J/A+A/365/409 J/A+A/413/843 Sohn S.T.: J/AJ/131/866 J/ApJ/623/181 Sohn Y.-J.: J/A+A/437/61 Sohn Y.J.: J/A+A/445/69 Sohn Y.-J.: J/A+A/454/717 J/AJ/112/2559 Sohn Y.J.: J/AJ/115/130 Sohn Y.-J.: J/AJ/126/2840 J/AJ/127/958

Soifer B.T.: J/AJ/110/1993 J/AJ/120/2331 J/AJ/121/97 J/AJ/125/2411 J/AJ/126/1607 J/AJ/131/250

J/ApJS/73/359 SOIFER B.T.: VII/109 Soifer T.: J/A+A/424/371 Soive A.: II/243 Soker N.: J/ApJS/112/487 Sokol G.V.: J/PAZh/25/30 J/PAZh/26/705 J/PAZh/27/187 Sokoloski J.L.: J/ApJ/636/1002 SOKOLOV N.A.: J/A+AS/110/553 Sokolov N.A.: J/A+AS/130/215 Sol H.: J/other/Ap/43.413 Solanes J.M.: J/A+AS/131/221 SOLANKI S.K.: J/A+AS/113/71 Solano E.: J/A+A/330/515 J/A+A/331/633 J/A+A/365/110 J/A+A/378/116 J/A+A/379/564 J/A+A/475/1159

J/A+AS/122/131 J/A+AS/125/321 Solarz M.: J/other/APh/26.282 SOLER M.: J/A+AS/110/351 Solivella G.: J/A+A/432/491 J/A+A/466/931 J/A+AS/118/231 J/MNRAS/367/1441 Solivella G.R.: J/other/NewA/12.461 J/other/RMxAA/39.89 Sollerman J.: J/A+A/428/555 J/A+A/429/559 J/AJ/130/2453 J/AJ/131/1648 J/ApJ/594/1 J/ApJ/602/571 Sollima A.: J/A+A/420/173 J/ApJ/640/L43 J/ApJ/654/915 J/MNRAS/354/708 J/MNRAS/357/265 SOLLNER T.C.L.G.: VII/50 Solomon A.R.: J/AJ/133/439 Solomon P.M.: VII/114 SOMA M.: I/188 Somerville R.: J/AJ/127/131 Somerville R.S.: J/ApJ/527/199 Sommer M.: J/ApJS/94/551 SOMMER-LARSEN J.: J/A+AS/109/171 Sommer-Larsen J.: J/AJ/117/2329 J/ApJS/96/175 Song C.: J/other/APh/26.282 Songaila A.: J/AJ/127/3121 J/AJ/127/3137 J/ApJ/471/L5 J/ApJ/538/29 J/ApJS/94/461 Sonneborn G.: J/ApJS/146/1 J/ApJS/146/125 J/ApJS/99/223 SOON W.H.: J/ApJ/416/787 Sopher J.E.: J/other/APh/26.282 Sordo R.: J/A+A/434/1107 Soria R.: J/A+A/410/53 Sosa Molina P.J.: J/MNRAS/341/1066 Sosey M.L.: J/ApJ/633/871 Sosin C.: J/A+A/391/945 J/AJ/109/639 Sostero G.: J/A+A/323/113 J/A+AS/116/1 Soszinski I.: II/278 J/AcA/51/317 J/AcA/52/129 J/AcA/54/347

Soszynski I.: J/A+A/450/873 J/AcA/48/563 J/AcA/49/223 J/AcA/49/319 J/AcA/49/437 J/AcA/50/1 J/AcA/51/175 J/AcA/51/221 J/AcA/52/143 J/AcA/52/217 J/AcA/53/1 J/AcA/53/93 J/AJ/128/1167 J/AJ/128/2815 J/ApJ/636/240 J/MNRAS/348/1439 J/MNRAS/356/331

Sota A.: J/A+A/449/937 Sotnikova Yu.V.: J/AZh/84/387 Soto M.: J/A+A/466/157 SOTO R.: J/A+AS/110/351 Soubiran C.: I/272 I/303 III/200B III/221A III/232 Soubiran C: III/251

Soubiran C.: J/A+A/373/576 J/A+A/398/141 J/A+A/411/559 J/A+A/418/551 J/A+A/427/933 J/A+A/438/139 J/A+AS/133/221 J/AZh/80/458 J/AZh/84/429 J/AZh/86/158 J/MNRAS/359/865 J/MNRAS/371/879 Soubiran: III/218 Soucail G.: J/A+A/363/440 J/A+A/372/391 J/A+A/421/509 J/A+A/468/937 J/A+AS/124/283 Soucail: J/A+A/394/395 SOUCHAY J.: J/A+A/299/696 Souchay J.: J/A+A/312/1017 J/A+AS/116/473 J/A+AS/125/399 Soundararajaperumal S.: J/ApJS/115/185 SOURIAU J.M.: VII/32 Sousa S.G.: J/A+A/458/873 J/A+A/469/783 Southwell K.A.: J/MNRAS/298/692 Southworth J.: J/A+A/456/651 J/A+A/461/1065 J/A+A/461/1077 J/A+A/467/1215 J/MNRAS/363/529

J/MNRAS/373/687 Sowards-Emmerd D.: J/AJ/119/2598 J/ApJ/590/109 J/ApJ/609/564 J/ApJ/626/95 Sowell J.R.: J/AJ/106/637 J/AJ/128/1319 J/AJ/129/2798 J/PASP/105/36 Sowers J.W.: J/ApJ/525/420 Soydugan E.: J/A+A/387/240 J/A+A/417/263 J/MNRAS/370/2013 J/MNRAS/373/435 Soydugan F.: J/A+A/387/240 J/A+A/417/263 J/MNRAS/370/2013 J/MNRAS/373/435 Sozzetti A.: J/AJ/131/2274 Space Telescope Science Institute: II/258 Spadafora A.L.: J/A+A/437/789 Spagna A.: J/A+A/341/L79 J/A+A/448/579 J/A+AS/130/359 J/MNRAS/375/1381 Spahn E.Y.: J/AJ/123/3409 Spahr T.: J/ApJ/626/L5 Spahr T.B.: J/AJ/131/527 Spalding R.: J/PASP/114/1382 Spangenberg E.E.: J/AZh/62/229 Spangler C.: J/ApJ/555/932 Sparke S.: J/A+A/386/140 Sparks P.: J/MNRAS/367/1609 Sparks W. B.: J/ApJS/122/81 Sparks W.B.: J/A+A/288/413 J/A+AS/138/253 J/AJ/128/1990 J/ApJ/513/733 J/ApJ/612/202 J/ApJ/621/53

J/ApJS/150/1 Spasojevic Dj.: J/A+A/477/967 Spassova N.: J/A+AS/110/1 J/A+AS/116/447 J/A+AS/121/499 Speck A.K.: J/A+AS/146/437 J/MNRAS/371/1744 Spekkens K.: J/AJ/129/2119 J/AJ/130/2613 Spellman K.: J/PASP/101/360 SPENCER J.H.: I/5 Spencer J.H.: I/6 Spencer Jones J.H.: J/other/SAAOC/15.65 J/other/SAAOC/15.73 SPERANDIO M.: J/A+AS/110/279 Sperandio M.: J/A+AS/137/101 Sperauskas J.: J/A+A/397/997 J/A+A/461/509 J/AN/323/139 J/BaltA/11/465 J/BaltA/8/325 Spergel D.N.: J/ApJS/148/97 Speroni N.: J/A+AS/103/541 Spezzi L.: J/A+A/470/281 Spiczak G.M.: J/other/APh/26.282 Spiekermann G.: J/A+AS/125/71 Spielfiedel A.: J/A+A/451/1125 Spiering C.: J/other/APh/26.282 Spinoglio L.: J/AJ/127/3075 J/ApJ/453/616 VII/157

Spinrad H.: II/47 J/AJ/112/1197 J/AJ/122/598 J/AJ/123/2223 J/AJ/130/1324 J/ApJ/600/L99

J/ApJ/613/200 Spinrad. H: J/PASP/97/932 Spiridonova O.I.: J/A+A/376/775 J/AZh/81/925 J/PAZh/27/83 J/PAZh/27/942 Spiro M.: J/A+A/303/137 J/A+A/389/149 J/A+A/454/185 SPIRO M.: J/A+AS/109/447 Spite F.: J/A+A/364/674 J/A+A/403/1105 J/A+A/416/1117 SPITE M.: J/A+A/293/347 Spite M.: J/A+A/364/674 J/A+A/403/1105 J/A+A/404/661 J/A+A/416/1117 Spitler L.: J/MNRAS/366/1230 Spitzak J.G.: J/ApJS/119/159 Spolaor M.: J/A+A/449/461 J/A+A/467/37 J/A+A/469/861 Spoon H.W.W.: J/A+A/390/501 J/A+A/390/511 J/A+A/418/465 J/A+AS/106/141 Sposetti S.: J/A+A/465/331 SPRAYBERRY D.: J/AJ/106/530 Sprayberry D.: J/ApJS/105/209 Spreeuw H.: J/A+A/475/359 Sprimont P.G.: J/A+A/474/473 Sprimont P.-G.: J/MNRAS/382/279 Springob C.M.: J/AJ/130/2613 VIII/77 Spyromilio J.: J/A+A/428/555 J/AJ/130/2453 J/AJ/131/1648 J/ApJ/594/1 J/ApJ/602/571 Squires G.: J/A+A/432/381 Squires G.K.: J/AJ/131/250 Sramek R.: J/AJ/80/771 Sramek R.A.: J/AJ/104/704 J/AJ/108/1163 Sreckovic A.: J/A+AS/102/607 Sreekumar P.: J/ApJS/101/259 J/ApJS/107/227 J/ApJS/123/79 J/ApJS/94/551 Srianand R.: J/A+A/473/791 Sridharan T.K.: J/A+A/417/115 J/ApJ/606/929 Srinivasan G.: J/other/JApA/25.143 Srividya S.: J/MNRAS/372/1315 STABELL R.: J/A+A/300/323 Stacey G.: J/AJ/129/2074 Stacy G.: J/A+AS/143/145 Stagg C.R.: J/AJ/109/359 Stagni R.: J/A+A/428/67 Stahl O.: J/A+A/304/415 J/A+A/389/931 J/A+A/398/49 J/A+A/418/885 J/A+A/441/181 J/A+AS/106/141

J/A+AS/133/299 Stahler S.W.: J/ApJ/581/1194 Stairs I.H.: J/MNRAS/328/17 J/MNRAS/335/275 J/MNRAS/342/1299 J/MNRAS/352/1439 J/MNRAS/359/1524

J/MNRAS/372/777 Stalin C.S.: J/MNRAS/351/1071 Stamatikos M.: J/other/APh/26.282 Stamm W.: J/A+A/390/39 J/A+A/406/L9 J/A+A/410/813 Stanek K.: J/A+AS/122/471 Stanek K.Z.: J/A+AS/128/19 J/AcA/48/455 J/AJ/115/1016 J/AJ/115/1894 J/AJ/117/2810 J/AJ/118/2211

J/AJ/118/346 J/AJ/121/2032 J/AJ/121/861 J/AJ/121/870 J/AJ/122/1383 J/AJ/122/2477 J/AJ/123/3460 J/AJ/126/175 J/AJ/127/861 J/AJ/128/1761 J/AJ/129/1596 J/AJ/129/2856 J/AJ/131/527 J/ApJ/591/L111 J/ApJ/626/L5 J/MNRAS/371/1405

Stanev T.: J/other/APh/26.282 Staneva A.: J/A+AS/116/447 Stanford S.A.: J/A+A/432/381 J/A+A/461/823 J/AJ/123/2223 J/AJ/123/637 J/AJ/127/131 J/AJ/127/2484

J/ApJ/504/113 J/ApJ/639/816 J/ApJS/131/185 J/ApJS/142/153 Stanga R. M.: J/ApJS/125/349

Stanghellini C.: J/A+A/363/887 J/A+A/369/380 J/A+A/379/755 J/A+A/426/463 J/A+A/450/959

J/A+AS/131/303 Stanghellini L.: J/A+AS/94/399 J/ApJ/622/294 Stanimirovic S.: J/MNRAS/339/105 Stanishev V.: J/A+A/437/789 J/A+A/456/599 Stanke T.: J/A+A/371/833 J/A+A/392/239 J/A+A/409/147 J/A+A/426/171 J/A+A/447/609 Stankov A.: J/ApJS/158/193 Stansberry J.A.: J/ApJ/620/1010 J/ApJ/636/1098 Stapelfeld K. R.: J/PASP/108/682 Stapelfeldt K.: J/AJ/113/1001 Stapelfeldt K.R.: J/AJ/117/1700 J/ApJ/620/1010 J/ApJ/636/1098 Stappers B.W.: J/A+A/445/243 J/A+A/469/607 Starck J.L.: J/A+A/342/313 Starck J.-L.: J/MNRAS/382/279 Stark A.A.: J/ApJ/517/282 STARK A.A.: VII/50 VII/76 Stark A.A.: VIII/10 VIII/28 Stark M.: J/AJ/123/3216 Stark M.A.: J/AJ/126/1455 STARK R.: J/A+AS/93/121 Starkey D.: J/other/IBVS/5526 Starkman L.: J/AJ/121/2331 STARON R.T.: III/11 III/12 Starrfield S.: J/A+A/434/1107 STASINSKA G.: J/A+A/303/893

Stasinska G.: J/A+A/318/256 J/A+A/327/736 J/A+A/336/667 J/A+A/352/297 J/A+A/427/231

J/A+A/441/981 J/A+A/448/955 J/A+A/466/75 J/A+A/469/799 J/A+AS/106/559 Stassun K.G.: J/A+A/382/899 J/AJ/117/2941 Stathopoulou M.: J/A+A/274/555 Statler T.S.: J/AJ/108/111 Stauber P.: J/A+A/475/549

Staubert R.: J/A+A/311/384 J/A+A/342/736 J/A+A/395/129 J/A+A/407/1039 J/A+A/466/41

J/A+AS/114/465 Staude H.J.: J/A+AS/127/1 Stauffer J.: J/A+A/341/163 J/A+A/341/751 J/AJ/127/2659 J/AJ/128/787 J/ApJ/638/897 J/PASP/118/1690

Stauffer J.R.: J/A+A/310/879 J/A+A/372/862 J/A+A/395/813 J/A+A/400/891 J/A+A/436/853

J/A+A/471/499 J/AJ/106/1059 STAUFFER J.R.: J/AJ/106/1080 Stauffer J.R.: J/AJ/110/1229 J/AJ/124/1570 J/AJ/124/601 J/AJ/131/527 J/ApJ/445/280 J/ApJ/546/1006

J/ApJ/593/1093 J/ApJ/610/1064 J/ApJ/635/560 J/ApJS/134/103 STAUFFER J.R.: J/ApJS/85/315 Stauffer J.R.: J/ApJS/91/625 Staveley-Smith L.: J/AJ/123/873 J/AJ/124/1954 J/AJ/124/690 Staveley-smith L.: J/AJ/128/16 Staveley-Smith L.: J/AJ/129/220 J/ApJS/165/307 J/MNRAS/289/225 J/MNRAS/335/1085 J/MNRAS/337/641

J/MNRAS/339/105 Staveley-smith L.: J/MNRAS/354/305 Staveley-Smith L.: J/MNRAS/355/44 J/MNRAS/371/898 Stavrev K.Y.: VI/90 Stawikowski A.: III/226 Stebbins J.: II/8 Stecher T.P.: J/AJ/108/1350 J/AJ/112/601 J/AJ/113/1011 J/AJ/116/180 J/AJ/121/891 STECHER T.P.: J/ApJ/413/604 J/ApJ/425/122 Stecher T.P.: J/ApJ/438/181 J/ApJ/446/622 J/ApJ/469/629 J/ApJS/104/287 J/ApJS/98/595 VI/104 Stecklum B.: J/A+A/434/987 J/ApJS/155/123 Stedman S.: J/A+A/414/23 VI/112 Stee P.: J/A+A/312/879 Steele D.: J/other/APh/26.282 Steele I.A: J/A+A/414/23 STEELE I.A.: J/A+AS/109/29 Steele I.A.: J/MNRAS/342/1241 Steemers W.J.G.: J/A+AS/73/103 Stefanik R.P.: J/A+AS/126/21 J/AJ/119/1389 J/AJ/124/1144 J/AJ/131/527 J/PASP/118/1656 Stefanil R.P.: J/A+A/399/105 Steffen A.T.: J/AJ/128/1483 J/AJ/128/1501 Steffen M.: J/A+AS/126/39 Steffen P.: J/A+AS/58/197 J/other/APh/26.282 Stefka V.: J/A+A/463/783 STEFL S.: J/A+A/294/135 Stefl S.: J/A+A/304/415 J/A+A/309/521 J/A+A/309/787 J/A+A/459/137 J/A+AS/125/75 Stefnik R.P.: J/ApJ/466/415 Stehle C.: VI/82 Stehle M.: J/MNRAS/369/1880 Stehle R.: J/A+A/312/93 Steidel C.C.: J/ApJ/480/568 J/ApJ/519/1 J/ApJ/592/728 J/ApJ/607/665 J/ApJ/613/200 J/ApJS/127/1

J/MNRAS/356/568 J/MNRAS/371/221 Stein P.: J/A+A/327/952 J/A+AS/116/203 J/A+AS/131/221 STEIN W.L.: V/21 Steinacker J.: J/A+A/474/883 Steindling S.: J/ApJ/621/53 Steine M.: J/A+A/447/921 Steiner J.E.: J/A+A/444/895 J/A+AS/124/55 J/A+AS/131/119 STEINER J.E.: J/MNRAS/259/233 Steinle H.: J/A+AS/143/145 Stella L.: J/A+A/458/21

Stelzer B.: J/A+A/342/736 J/A+A/353/1044 J/A+A/359/227 J/A+A/423/1029 J/A+A/439/1149

J/A+A/468/353 Stenholm B.: J/A+A/327/191 J/A+A/331/949 J/A+AS/106/559 V/100 V/84 Stepanian D.A.: VII/61A Stepanian J.A.: J/A+AS/147/169 J/AJ/124/1283 STEPANIAN J.A.: VII/172 Stepanova E.A.: I/44 Stephen J.B.: J/ApJ/607/L33 J/ApJ/636/765 J/ApJS/170/175 Stephens A.W.: J/A+A/474/121 J/AJ/132/271 Stephens R.: J/A+A/465/331 Stephenson C. B.: J/ApJS/86/453 Stephenson C.B.: III/123B III/15 III/15 III/15 III/156 III/180 STEPHENSON C.B.: III/36 Stephenson C.B.: III/43 III/60 III/76A III/76A III/76A J/AJ/80/972 J/ApJ/300/779 J/ApJ/301/927 STEPHENSON: III/168 Steppe H.: J/A+A/85/329 J/A+AS/102/611 STERKEN C.: II/188 II/200 Sterken C.: II/202 J/A+A/304/415 J/A+A/312/879 J/A+A/343/847 J/A+A/374/1056 J/A+A/389/931 Sterken C: J/A+A/418/249 Sterken C.: J/A+A/456/599 J/A+A/463/1053 J/A+AS/106/141 J/A+AS/121/113 J/AJ/120/2101 Stern B.: J/ApJS/156/217 Stern B.E.: J/ApJ/563/80

Stern D.: J/A+A/434/53 J/A+A/454/423 J/A+A/478/83 J/AJ/122/598 J/AJ/123/2223 J/AJ/123/637 J/AJ/127/180 J/AJ/130/1324 J/ApJ/596/944 J/ApJ/600/L99 J/ApJ/613/200 J/ApJ/639/816 J/ApJS/131/185 J/ApJS/156/35 J/ApJS/165/19

Stern R.A.: J/AJ/110/1229 J/ApJ/325/798 J/ApJ/348/253 J/ApJ/448/683 Sternberg A.: J/AJ/121/3048 J/ApJ/566/880 J/ApJ/643/1011 Sternberg J.R.: VI/43 Sterzik M.: J/A+A/326/211 J/A+A/333/619 J/A+A/450/681 J/A+A/460/695

Sterzik M.F.: J/A+A/297/391 J/A+A/312/439 J/A+A/325/647 J/A+A/384/1030 J/A+A/433/151

J/A+AS/124/449 J/A+AS/135/319 J/A+AS/146/323 Stetson B.: J/PASP/99/739 Stetson P.B.: J/A+A/471/515 J/A+AS/131/89 J/AJ/113/1001 J/AJ/113/1026 J/AJ/117/247 J/AJ/127/1579

J/AJ/127/2771 J/AJ/130/1177 J/AJ/97/1360 J/ApJ/463/26 J/ApJ/507/655 J/ApJ/508/491 J/ApJ/512/48 J/ApJ/514/614 J/ApJ/521/155 J/ApJ/523/540 J/ApJ/527/199 J/ApJ/529/723 J/ApJ/627/579 J/ApJS/120/265 J/MNRAS/371/1405 J/MNRAS/378/318 J/PASP/109/883 J/PASP/110/1172 J/PASP/110/533 J/PASP/115/413 J/PASP/116/1012 J/PASP/117/1325

Stevens H.L.: J/ApJ/484/394 Stevens I.R.: J/A+A/395/499 Stevens J.A.: J/AJ/130/1418 J/MNRAS/372/1621 Stevenson M.: J/ApJ/627/621 Stevenson S.L.: J/AJ/124/3465 Stewart B.C.: J/MNRAS/274/1165 Stewart D.I.: J/A+AS/65/485 Stewart G.C.: J/A+A/476/1191 J/MNRAS/349/135 J/MNRAS/352/91 Stewart I.: J/AJ/123/873 J/AJ/129/220 Stewart I.M.: J/AJ/124/1954 J/AJ/124/690 J/AJ/128/16 J/MNRAS/346/787 J/MNRAS/355/1303 Stewart S.G.: J/AJ/120/1794 Stezelberger T.: J/other/APh/26.282 Stiavelli M.: J/AJ/120/2747 Stickel H.: J/A+AS/115/11 Stickel: J/A+A/466/1205

Stickel M.: III/175 J/A+A/359/865 J/A+A/372/364 J/A+A/401/873 J/A+A/422/39 J/A+A/453/923

J/A+AS/105/211 J/A+AS/105/67 J/A+AS/115/1 J/A+AS/98/393 Stiele H.: J/A+A/465/375 Stiening R.: II/246 J/AJ/125/2521 Stierwalt S.: J/AJ/130/2613 Stift M.J.: J/A+A/374/265 Still J.M.: II/202 Stinson G.: J/AJ/126/2209 Stirling A.M.: J/AJ/130/1418 Stirpe G. M.: J/ApJ/510/659 Stirpe G.M.: J/A+A/339/34 J/A+A/342/1 J/A+AS/130/323 J/ApJS/115/185 St-Louis N.: J/A+A/347/583 Stobbart A.-M.: J/MNRAS/370/25 Stobie E.: J/AJ/117/17 J/AJ/130/1 J/PASP/118/1690 Stobie R.S.: J/MNRAS/238/709 J/MNRAS/287/867 J/MNRAS/320/451 STOCK J.: I/216A III/14 Stock J.: III/15 STOCK J.: III/172

Stock J.: III/76A J/A+A/18/341 J/A+AS/24/35 J/AJ/127/1158 J/AJ/89/1897 J/other/POAN/2.59

J/other/RMxAA/33.83 J/other/RMxAA/33.83 J/other/RMxAA/38.129 J/PASP/110/1434 Stock M.: J/other/RMxAA/38.129 Stocke J.T.: IX/15 J/ApJ/566/744 J/ApJ/569/689 Stoecklin T.: J/A+A/430/1139 Stok J.: III/15 III/76A Stokstad R.G.: J/other/APh/26.282 Stoll D.: VII/196 Stolovy S.R.: J/ApJ/635/560 Stolyarov V.A.: J/A+A/370/78 J/other/BSAO/46.28 Stone R. G.: J/A+AS/119/489 Stone R.C.: I/263 J/AJ/114/850 J/AJ/115/296 J/AJ/119/2008 Stone R.P.S.: J/AJ/121/1872 J/ApJ/636/1002 J/ApJS/105/419 Stonkute R.: J/PASJ/56/1025 Stootman F.: J/AJ/123/873 J/AJ/124/1954 J/AJ/124/690 J/AJ/128/16 J/MNRAS/337/641 Storchi-Bergmann T.: J/ApJS/105/93

Storey P.J.: J/A+A/370/1071 J/A+A/387/1135 J/A+A/394/753 J/A+AS/131/153 J/A+AS/133/257

J/A+AS/135/159 J/A+AS/135/359 J/A+AS/137/157 J/A+AS/142/85 J/MNRAS/368/1959 VI/64 STORM J.: J/A+A/291/121 Storm J.: J/A+A/364/455 J/A+A/402/509 J/A+A/415/521 J/A+A/415/531 J/A+A/440/487 J/A+A/476/73

J/A+AS/122/559 J/AJ/112/2408 J/AJ/128/1167 Storrie-Lombardi L.J.: J/A+A/370/426 J/AJ/117/17 J/AJ/125/2411 J/AJ/128/1 J/AJ/131/250 J/ApJ/543/552

J/ApJ/546/694 Storrie-lombardi L.J.: J/MNRAS/346/1103 Storrie-Lombardi M.C.: J/MNRAS/274/1107 Stott J.P.: J/MNRAS/379/1343 Stoufer M.C.: J/other/APh/26.282 Stoughton C.: J/AJ/119/2571 J/AJ/121/2308 J/AJ/121/2331 J/AJ/123/567 J/AJ/126/2579 J/AJ/127/1811

J/AJ/127/2210 J/AJ/130/873 J/ApJS/122/355 J/MNRAS/372/425 Stoy R.H.: I/116 STOY R.H.: I/75 Stoy R.H.: II/222 Stoyanov S.: J/other/APh/26.282

Strader J.: J/A+A/443/413 J/AJ/131/814 J/ApJ/637/1102 J/ApJS/169/430 J/MNRAS/343/665

J/MNRAS/366/1230 STRAIZYS V.: II/130 II/157A II/196

Straizys V.: II/93 J/BaltA/10/319 J/BaltA/10/349 J/BaltA/11/219 J/BaltA/11/231 J/BaltA/11/321 J/BaltA/11/417 J/BaltA/12/301 J/BaltA/12/323 J/BaltA/13/1 J/BaltA/13/555 J/BaltA/15/483Short J/BaltA/7/369

Straniero O.: J/A+A/417/269 J/AJ/118/2839 J/ApJ/654/915 STRASBOURG: I/49 Strasser S.: J/A+A/458/569 J/A+A/475/1053 Strassmeier K. G.: J/A+AS/142/275

Strassmeier K.G.: J/A+A/362/673 J/A+A/374/1049 J/A+A/399/315 J/A+A/440/735 J/A+A/454/295

J/A+A/476/881 J/A+AS/125/11 J/A+AS/137/369 J/A+AS/140/29 J/AN/321/277 STRASSMEIER K.G.: V/71 Strassmeier K.G.: V/76 Strateva I.: J/AJ/126/2125 Strateva I.V.: J/AJ/126/1720 J/AJ/130/387 Stratta G.: J/A+A/474/793 Straubmeier C.: J/A+A/433/117 Straughn A.N.: J/ApJ/639/724 Straume J.: J/A+A/401/531 Straume J.I.: J/A+AS/118/105 STRAUSS M.: II/174 Strauss M.: VII/185 VII/185 Strauss M.A.: J/AJ/121/2308 J/AJ/121/31 J/AJ/122/503 J/AJ/123/3409 J/AJ/123/567 J/AJ/124/1651 STrauss M.A.: J/AJ/126/1023 Strauss M.A.: J/AJ/126/1720 J/AJ/126/2125 J/AJ/126/2209 J/AJ/126/2579 J/AJ/127/3553 J/AJ/128/1002

J/AJ/129/2542 J/AJ/130/873 J/AJ/131/1 J/AJ/131/49 J/ApJ/607/426 J/ApJS/109/333 J/ApJS/155/257 VII/185 VII/198

Straw S.M.: J/ApJS/69/99 Streblyanska A.: J/A+A/444/79 J/A+A/451/457 Street R.A.: J/A+A/467/785 J/MNRAS/358/795 J/MNRAS/362/117 J/MNRAS/372/1117 J/MNRAS/379/647

J/MNRAS/379/816 Strel’nikov V.V.: J/PAZh/22/269 Strickland D.K.: J/ApJ/602/231 Strigachev A.: J/A+A/422/1023 J/A+A/464/641 STROBEL A.: III/121 Strobel A.: J/ApJ/634/451 STROBEL J.: III/121 Strobel K.: III/116 Strolger L.-G.: J/ApJ/607/665 J/ApJ/613/200 STROM C.: J/AJ/107/1453 Strom K.M.: J/AJ/109/1379 STROM K.M.: J/ApJ/438/813 Strom S.: J/AJ/127/2659 J/AJ/128/787 J/PASP/118/1690 Strom S.E.: III/240 J/AJ/106/1906 STROM S.E.: J/AJ/106/656 Strom S.E.: J/AJ/114/198 J/AJ/116/1816 J/AJ/119/3026 J/AJ/123/1528 J/AJ/127/2228 J/AJ/129/809

J/AJ/132/271 STROM S.E.: J/ApJ/438/813 Strom S.E.: J/ApJ/601/979 Strong A.: J/A+AS/143/145 Strong A.W.: J/MNRAS/201/495 Strong S.B.: J/AJ/131/2478 Stroozas B.A.: J/AJ/107/751 Struble M.F.: J/AJ/108/1 J/ApJS/125/35 VII/177 STRUBLE M.F.: VII/56 Struble M.F.: VII/95 STRUBLE M.F.: VII/96 Struck C.: J/AJ/121/182 Stryker L.L.: III/88 J/ApJ/298/544 STScI Team: B/hst Stubbs C.: J/AJ/121/2308 J/AJ/121/2331 J/ApJ/542/257 J/ApJ/598/597 Stubbs C.W.: J/A+A/415/863 J/AJ/114/326 J/AJ/130/2453 J/AJ/131/1648 J/ApJ/562/337 J/ApJ/594/1

J/ApJ/602/571 Studer M.: J/A+A/324/137 Studt J.: J/A+AS/127/145 Stuetz C.: J/A+A/441/631 STUEWE J.A.: VII/68A Sturch C.: J/AJ/112/2408 Sturch C.R. et al.: II/143A Sturmann L.: J/A+A/453/155 Sturmann2 J.: J/A+A/453/155 Sturner S.: J/ApJS/156/69 Stutz A.: J/ApJS/143/73 Stutzki J.: VIII/66 VIII/66 Su C.G.: J/A+AS/118/503 Su C.-G.: J/A+AS/128/255 Su H.: J/AJ/117/2757 Su K.Y.L.: J/ApJ/620/1010 J/ApJ/636/1098 Suarez J.C.: J/A+A/450/715 J/A+A/456/261 J/A+A/472/565 Suarez O.: J/A+A/458/173 Subasavage J.P.: J/AJ/129/413 J/AJ/130/1658 Subbarao M.: J/AJ/121/2308 SubbaRao M.: J/AJ/121/2331 Subbarao M.: J/AJ/123/567 J/AJ/125/1817 J/AJ/126/1023 J/AJ/126/2579 J/AJ/131/2018 J/ApJ/588/824

J/PASJ/55/771 Subrahmanya C.R.: J/ApJS/107/19 Subrahmanyan R.: J/AJ/130/896 J/MNRAS/354/305 J/MNRAS/360/1305 J/MNRAS/371/898 Subramaniam A.: J/A+AS/127/167 J/A+AS/135/391 J/A+AS/145/365 J/MNRAS/371/1301 Subramaniam: J/A+A/430/421 Suchkov A.A.: J/AJ/130/2439 J/ApJ/595/1206 J/ApJ/612/437 J/ApJ/616/1284 J/ApJS/147/103 Sudar D.: J/A+A/350/566 J/A+A/387/580 J/A+A/408/611 J/A+A/415/289 J/A+A/416/669 J/A+A/455/1037

J/A+A/463/1061 J/A+A/463/233 J/A+A/464/263 Sudzius J.: J/BaltA/7/369 Sugai H.: J/ApJS/142/79 Suganuma M.: J/ApJS/141/23 J/ApJS/141/31 Sugar J.: VI/73 VI/74 Sugars B.A.: J/AJ/109/359

Sugitani K.: J/A+A/309/581 J/AJ/123/2597 J/ApJ/596/1064 J/MNRAS/354/980 J/MNRAS/368/534

J/PASJ/55/1079 J/PASJ/58/L29 Sugiyama N.: J/ApJS/100/281 Sugizaki M.: J/ApJS/134/77 Suhhonenko I.: J/A+A/459/L1 J/A+A/462/811 J/AN/327/365 Sujatha S.: J/other/BASI/32.295 Sulanke K.-H.: J/other/APh/26.282 Suleiman R.: J/A+AS/143/145 Sulentic J.: J/A+A/436/443 J/A+A/462/507 J/A+A/470/505 J/A+A/472/121 SULENTIC J.W.: J/A+A/297/28 Sulentic J.W.: J/A+A/449/937 J/A+AS/110/131 J/ApJS/145/199 VII/1B Sullivan D.J.: J/MNRAS/330/137 J/MNRAS/331/609 Sullivan G.W.: J/other/APh/26.282 Sullivan M.: J/ApJ/624/135 J/ApJS/155/1 J/MNRAS/312/442 J/MNRAS/330/75 SULLIVAN W.T.III: J/A+AS/113/35 Sumi T.: J/ApJ/636/240 J/MNRAS/330/137 J/MNRAS/331/609 J/MNRAS/340/1346 J/MNRAS/348/1439

J/MNRAS/356/331 J/MNRAS/370/435 Sumiya R.: J/PASJ/57/881 Sumner T.: J/MNRAS/351/1290 Sumner T.J.: J/other/APh/26.282 Sun J.: J/A+AS/127/181 J/other/AcApS/18.302 J/other/PPMtO/14.185 Sun W.: J/A+A/397/361 Sun W-H.: J/AJ/112/628 Sun W.-H.: J/AJ/117/2757 J/AJ/126/1286 Sun Y.-C.: J/other/AcApS/18.302 SUNDMAN A.: III/33 Sundman A.: III/35

Sung H.: J/A+A/454/1047 J/AJ/114/2644 J/AJ/115/734 J/AJ/120/333 J/AJ/123/290 J/AJ/127/1014

J/AJ/128/1684 J/MNRAS/306/361 J/MNRAS/310/982 Suntzeff N.B.: II/179 J/A+A/415/863 J/AJ/108/1722 J/AJ/110/2747 J/AJ/112/2408 J/AJ/113/634

J/AJ/127/2674 J/AJ/128/1558 J/AJ/130/2453 J/AJ/131/1648 J/ApJ/562/303 J/ApJ/594/1 J/ApJ/602/571 J/MNRAS/300/665

Sunyaev R.: J/A+A/474/793 J/A+A/475/775 J/A+AS/130/233

Sunyaev R.A.: IX/36 J/A+AS/129/1 J/PAZh/26/355 J/PAZh/28/323 J/PAZh/30/430 J/PAZh/30/589

J/PAZh/32/508 SuperCOSMOS Sky Surveys: VI/124 Supper R.: J/A+A/278/36 J/A+A/317/328 J/A+A/359/227 J/A+A/373/63 J/A+AS/114/215 Surace C.: J/A+AS/133/171 J/MNRAS/351/1290 Surace J.: J/A+A/467/565 J/AJ/129/1198 J/AJ/129/2074 J/MNRAS/371/1891 Surace J.A.: II/255 J/A+A/467/73 J/AJ/126/1607 J/AJ/127/3235 J/AJ/131/250 Suran M.D.: J/A+A/327/1017 J/A+A/348/184 Surdej J.: J/A+A/439/413 J/A+A/467/73 J/A+A/474/473 J/A+AS/123/529 J/MNRAS/382/279 Surkova L.P.: V/115 Surma P.: J/ApJS/107/1 Suso J.: J/A+A/472/565 J/A+AS/115/315 J/A+AS/123/1 J/A+AS/123/59 Suso Lopez J.: J/A+A/304/415 Sutantyo W.: J/A+A/331/949 Sutherland R.S.: J/ApJ/478/603 J/ApJ/619/755 Sutherland W.: J/AJ/114/326 J/AJ/127/334 J/ApJ/542/257 J/ApJ/552/289 J/ApJ/562/337 J/ApJ/598/597

J/MNRAS/329/227 J/MNRAS/329/87 J/MNRAS/333/100 J/MNRAS/350/1485 J/MNRAS/352/939 Sutherland W.J.: J/MNRAS/246/433 J/MNRAS/279/459 J/MNRAS/289/263 Sutherland W.J. Mon. Not. R. Astron. Soc. 269: J/MNRAS/269/151 Sutherland W.J.: VII/221 Suto H.: J/A+A/449/583 Sutorius E.: J/A+A/398/49 Sutton J.M.: VIII/2 SUZIKI S.: I/167 SUZUKI S.: I/186 I/187 I/188 I/198 Suzuki S.: I/248 Suzuyama T.: J/PASJ/52/631 Svatek G.F.: VI/34 Svechnikov M.A.: J/A+A/437/375 V/115 V/118 V/121 V/124 V/42 Svensson R.: J/ApJ/563/80 Sviderskiene Z.: VI/50 Swank J.H.: J/ApJS/96/303 Swartz D.A.: J/ApJ/615/897 J/ApJS/144/213 J/ApJS/154/519 Swaters R.: J/A+A/420/147 Swaters R.A.: J/A+A/390/863 J/A+A/442/137 SWEENEY L.H.: II/161 Sweeney L.H.: II/96 Sweigart A.V.: J/AJ/119/1424 J/AJ/119/1448 J/AJ/127/2738 J/AJ/130/1693 J/ApJ/532/308 J/ApJ/562/368 Swensson A.: J/AJ/132/137 Swift C.: III/214 Swifte R.H.D.: I/144 Swinbank A.M.: J/MNRAS/379/1343 Swings J.-P.: J/A+AS/119/1 Swinyard B.M.: J/A+A/406/L47 Sykes C.M.: VIII/72 Sykes M.: II/246 Sylvester R.J.: J/A+AS/146/437 Syphers D.: J/AJ/124/2039 Szabados L.: J/A+A/461/613 Szabo G.M.: J/A+A/375/285 J/A+A/376/561 J/A+A/410/879 Szabo Gy.M.: J/A+A/374/712 J/other/P+SS/53.925 Szaidl B.: J/A+A/469/1033 Szalai T.: J/A+A/465/943 Szalay A.: J/AJ/121/2331 J/ApJS/159/41 Szalay A. S.: J/ApJS/122/355

Szalay A.S.: J/AJ/121/2308 J/AJ/121/31 J/AJ/123/567 J/AJ/126/2579 J/AJ/127/1811 J/AJ/130/825

J/AJ/131/84 J/ApJ/619/L95 J/ApJS/164/38 VII/200 Szatmary K.: J/A+A/390/585 J/A+A/394/97 J/A+A/463/589 Szchwarzenberg-Czerny A.: J/A+A/424/1101

Szczerba R.: J/A+A/352/297 J/A+A/368/1006 J/A+A/434/201 J/A+A/469/799 J/A+AS/126/39

J/PAZh/26/115 J/PAZh/26/510 SZCZODROWSKA-KOZAR B.: J/A+AS/110/411 Szczygiel D.: II/264 Szeidl B.: J/A+A/430/1049 J/A+A/445/617 SZEIFERT T.: J/A+A/289/740 Szeifert T.: J/A+A/389/931 J/A+A/398/49 J/A+A/459/423 Szekely P.: J/A+A/394/943 J/A+A/463/589 J/other/P+SS/53.925 Szentgyorgyi A.: J/AJ/129/1596

Szentgyorgyi A.H.: J/AJ/121/2032 J/AJ/122/2477 J/AJ/123/3460 J/AJ/127/861 J/AJ/129/2856

J/AJ/129/363 J/AJ/131/527 Szewczyk O.: II/278 J/AcA/51/221 J/AcA/51/317 J/AcA/52/143 J/AcA/52/217 J/AcA/53/1 J/AcA/53/93

J/AcA/54/347 J/ApJ/636/240 J/MNRAS/348/1439 J/MNRAS/356/331 Sziladi K.: J/A+A/376/561 Szing A.: J/A+A/430/1049 Szkody P.: J/A+A/444/643 J/AJ/112/2408 J/AJ/123/3409 J/AJ/126/1023 J/AJ/131/1674 J/ApJ/607/426

J/MNRAS/373/687 Szokoly G.: II/253 J/A+A/371/833 J/A+A/393/425 J/A+A/398/901 J/A+A/451/457 J/A+A/476/1191

J/ApJS/155/73 J/ApJS/161/21 Szokoly G.P.: J/AJ/121/2308 J/AJ/121/2331 J/AJ/122/503 J/AJ/123/1807 J/ApJS/155/271 Sztajer P.: VIII/82 Szymanski J.: J/AJ/127/2436 Szymanski J.A.: J/AJ/119/1901 Szymanski: J/A+AS/120/139 J/A+AS/122/471 Szymanski M: I/244A Szymanski M.: II/278 J/A+AS/112/407 J/A+AS/125/343 J/A+AS/128/19 J/AcA/48/147 J/AcA/48/175

J/AcA/48/563 J/AcA/49/223 J/AcA/49/319 J/AcA/49/437 J/AcA/49/521 J/AcA/50/1 J/AcA/51/175 J/AcA/51/221 J/AcA/51/259 J/AcA/51/317 J/AcA/52/129 J/AcA/52/143 J/AcA/52/217 J/AcA/53/1 J/AcA/53/93 J/AJ/110/1141 J/AJ/128/2815 J/ApJ/636/240 J/MNRAS/348/1439

Szymanski M.K.: J/AcA/54/347 J/AcA/56/253 J/MNRAS/356/331 SZYMCZAK M.: J/A+A/296/727 J/A+A/297/494 Szymczak M.: J/A+A/322/159 J/A+A/414/235 J/A+A/434/613 J/A+AS/143/269 Szymkowiak A.E.: J/ApJS/96/303 T.: J/A+A/385/1014 J/A+A/404/775 J/A+A/475/915 Taaffe L.: J/other/PASAu/10.310 Tabara H.: J/A+AS/39/379 Tabatabaei F.S.: J/A+A/472/785 Taboada I.: J/other/APh/26.282 Tacconi L.: J/ApJ/606/664 Tachiev G.I.: J/A+A/385/716 Tachihara K.: J/PASJ/53/1081 Tadjibaev I.: J/A+A/440/53 Tadros H.: VII/221 Tafalla M.: J/ApJ/526/788 Taff L.G.: I/176 I/202 Tagliaferri G.: IX/34 J/A+A/364/205 J/A+A/375/739 J/A+A/384/491 J/A+A/384/491 J/A+A/418/509

J/A+A/435/799 J/A+A/448/L9 J/A+A/469/503 J/A+A/474/793 J/A+AS/115/41 J/A+AS/138/87 Tago E.: J/A+A/336/35 J/A+A/459/L1 J/A+A/462/811 J/A+AS/123/119 J/AJ/122/2222 J/AJ/123/51

J/AN/327/365 J/MNRAS/269/301 TAJAHMADY F.: J/A+AS/109/447 Tajer M.: J/A+A/435/799 J/A+A/439/413 J/A+A/467/73 J/A+A/474/473 Tajhamady F.: J/A+A/303/137 Tajitsu A.: J/AJ/115/1989 Takaba H.: J/PASJ/53/517 Takada-Hidai M.: J/ApJ/472/509 Takada-hidai M.: J/PASJ/57/13 J/PASJ/58/389 Takagi S.: J/PASJ/55/635 Takagi T.: J/MNRAS/372/1621 Takahashi I.: J/ApJ/645/256 Takahashi K.: J/A+A/441/1195 Takahashi T.: J/ApJ/518/656 J/ApJS/133/1 J/ApJS/148/275 J/ApJS/161/185 Takalo L.O.: J/A+AS/124/559 J/A+AS/128/459 J/A+AS/129/577 J/A+AS/130/109 J/A+AS/143/357 Takami H.: J/A+AS/125/399 Takamiya M.: J/AJ/110/1083 J/AJ/121/935 J/ApJS/122/109 TAKAMIYA M.Y.: J/AJ/106/2575 TAKASE B.: VII/104 VII/135 VII/150 Takase B.: VII/179A TAKATA T.: J/A+A/299/347 Takata T.: J/A+AS/104/529 J/AJ/111/42 J/ApJS/112/245 J/ApJS/89/57 J/MNRAS/307/236 J/MNRAS/342/802

J/MNRAS/362/1371 J/MNRAS/372/741 J/PASJ/53/653 J/PASJ/55/1079 TAKATA T.: VII/140 Takeda A.: J/ApJS/89/57 Takeda Y.: J/A+A/402/343 J/A+A/420/673 J/A+A/468/663 J/PASJ/57/109 J/PASJ/57/13 J/PASJ/57/27

J/PASJ/57/45 J/PASJ/57/65 J/PASJ/58/389 J/PASJ/59/245 Takeuchi T.: J/AJ/111/42 Takeuchi T.T.: J/ApJS/112/245 J/ApJS/164/38 J/MNRAS/380/986 Takeuti M.: J/MNRAS/330/137 J/MNRAS/331/609 Talavera A.: J/A+AS/137/351 J/ApJS/99/223 TALBERT F.D.: J/PASP/103/536 Talbi D.: J/A+A/319/318 Tallon-Bosc I.: J/A+A/312/879 Talon R.: IX/36 Tamagawa T.: J/other/GCN/1070.1 Tamazian V.S.: J/A+A/366/868 J/A+AS/126/471 J/AJ/127/1181 J/AJ/131/2681 Tamblyn P.: J/AJ/99/1079 J/ApJS/90/83 Tamburelli F.: J/MNRAS/327/771 Tammann G.A.: J/A+A/404/423 J/AJ/90/1681 J/ApJ/522/802 J/ApJ/551/973 VII/51 Tamura H.: J/PASJ/57/165 Tamura: J/A+A/385/1014

Tamura M.: J/AJ/131/1608 J/ApJ/596/1064 J/ApJS/118/177 J/MNRAS/353/705 J/MNRAS/354/980

J/MNRAS/361/211 J/MNRAS/368/534 J/PASJ/55/1079 J/PASJ/58/L29 Tamura N.: J/PASJ/53/653 Tamura S.: J/A+A/367/199 J/AJ/115/1989 Tamura T.: J/A+A/420/135 Tamura Y.: J/PASJ/52/631 Tanabe T.: J/A+A/442/281 J/ApJS/155/401 J/MNRAS/353/705 J/MNRAS/370/1979 Tanaka I.: J/ApJ/528/123 J/ApJ/634/861 J/MNRAS/365/1392 Tanaka M.: J/A+A/376/997 J/A+A/403/943 J/MNRAS/365/1392 J/MNRAS/379/1546 J/PASP/117/199 Tananbaum H.: IX/3 J/ApJ/618/123 J/ApJ/644/829 J/ApJ/645/955 J/ApJS/92/53 TANANBAUM H.: VII/18 Tananbaum H.D.: IX/13 J/ApJS/150/19 Tancredi G.: J/A+AS/115/295 J/A+AS/146/73 Tandon S.N.: J/A+AS/141/175 Tang G.S.: J/A+AS/127/181 J/other/PPMtO/14.185 Tang X.: VII/168 Tang Y.: J/A+AS/124/259 Tang Z.H.: J/A+A/454/379 J/MNRAS/376/1707 Taniguchi Y.: J/A+A/365/409 J/A+A/413/843 J/ApJ/648/7 J/PASJ/57/165 J/PASJ/57/881 J/PASJ/58/313 Tankosic D.: J/A+A/399/795 Tanner A.M.: J/AJ/112/62 Tanner D.: J/AJ/131/461 Tanriverdi T.: J/A+A/415/283 J/A+A/420/1039 Tantalo R.: J/A+A/311/361 J/A+A/335/823 Tanvir N.: J/MNRAS/273/157 Tanvir N.R.: J/A+A/411/361 J/AJ/112/359 Tanzellanitti G.: VII/99 Tanzi E.G.: J/AJ/104/28 Tao J.: J/A+A/369/57 J/A+A/369/758 J/ApJS/115/185 Tapia M.: J/A+A/316/102 J/A+A/357/1020 J/A+A/445/971 J/MNRAS/367/513 Tapken C.: J/A+A/413/1037 J/A+A/418/885 Tappert C.: J/A+AS/121/247 J/MNRAS/306/279 Tappin S.J.: J/A+A/440/373 J/MNRAS/229/589 Taraji H.: I/152 I/153 TARAJI H.: I/77 TARANOVA O.G.: J/A+AS/93/383

Taranova O.G.: J/AZh/79/1118 J/PAZh/25/860 J/PAZh/26/695 J/PAZh/27/393 J/PAZh/30/605

J/PAZh/32/866 Tarasov A.E.: J/A+A/405/1087 Tarasova O.: J/other/APh/26.282 Tarasova T.N.: J/AZh/79/526 Tarchi A.: J/PASJ/57/135 TARENGHI M.: J/AJ/107/1629 Tarenghi M.: J/MNRAS/276/689 Taresch G.: J/A+A/321/531 Targett T.A.: J/MNRAS/373/1531 Taribello: J/A+A/421/937 Tarusawa: J/A+A/374/504 Tarusawa K.: J/ApJS/97/77 Tas G.: J/A+A/374/1056 J/A+A/388/298 J/MNRAS/376/573 Tasca L.: J/A+A/468/937 Tashiro M.: J/ApJ/645/256 Tasse C.: J/A+A/456/791 J/A+A/471/1105 J/MNRAS/382/279 Tautvaisiene G.: J/A+A/360/499 J/A+A/380/578 J/A+A/431/933 J/AJ/127/373 J/BaltA/13/555 Tavastherna K.N.: I/82 TAVASTSHERNA K.N.: I/43 Tavolganskaya N.S.: J/AZh/83/265 Tayal S.S.: J/A+A/418/363 J/A+A/426/717 J/ApJ/526/544 J/ApJS/123/295 J/ApJS/163/207 Taylor A.C.: J/MNRAS/341/1066 J/MNRAS/342/915 J/MNRAS/360/340 TAYLOR A.R.: J/A+A/290/609 Taylor A.R.: J/ApJ/616/110 J/ApJ/626/195 J/ApJS/107/239 J/ApJS/131/95 J/ApJS/145/213 J/ApJS/155/33 Taylor: B/cgps Taylor B.G.: J/A+AS/109/9 J/A+AS/134/287 Taylor B.J.: II/47 TAYLOR B.J.: III/27 III/28

Taylor B.J.: J/A+A/398/721 J/A+A/398/731 J/A+AS/134/523 J/AJ/132/111 J/ApJS/159/100

J/ApJS/161/444 J/PASP/106/452 V/88 Taylor D.B.: J/A+AS/121/65 J/A+AS/130/77 J/A+AS/136/257 Taylor D.K.: J/ApJ/315/104 Taylor E.: J/ApJ/613/200 Taylor E.N.: J/AJ/134/1103 Taylor G.B.: J/A+A/472/763 J/AJ/130/923 J/ApJ/589/733 J/ApJ/634/L89 J/ApJS/100/1 J/ApJS/95/345

J/ApJS/98/1 J/ApJS/98/33 VIII/65 Taylor H.M.: J/ApJ/586/1356 Taylor J.H.: VII/156 VII/189 TAYLOR K.: J/MNRAS/244/408 J/MNRAS/253/686 Taylor K.: J/MNRAS/329/227 J/MNRAS/329/87 J/MNRAS/333/100 J/MNRAS/350/1485 J/MNRAS/352/939 Taylor M.: J/ApJ/549/820 Taylor R.: J/A+AS/125/399 J/PASJ/52/997 Taylor R.N.: VII/128 Taylor S.C.: J/ApJ/549/441 Taylor V.A.: J/ApJ/630/784 te Lintel Hekkert P.: J/A+A/267/515 Te Lintel Hekkert P.: J/A+AS/119/459 te Lintel Hekkert P.: J/A+AS/90/327 Teague P.F.: J/A+AS/93/61 J/A+AS/97/587 J/ApJS/72/715 Teays T.: J/ApJS/99/223 Tedds J.A.: J/A+A/428/383 J/A+A/469/27 J/A+A/474/473 J/A+A/476/1191 Tedesco E.: V/107 Tedesco E.F. (ed.): II/190 Tedesco E.F.: J/AJ/123/2070 Teegarden B.J.: J/ApJS/156/69

Teerikorpi P.: J/A+A/353/77 J/A+A/372/276 J/A+A/384/1050 J/A+AS/102/57 J/A+AS/129/577

J/A+AS/130/333 J/A+AS/143/357 Tegler S.C.: J/ApJ/626/L5 Teixeira R.: I/303 J/A+A/361/1143 J/A+A/375/308 J/A+A/383/1054 J/A+A/409/361 J/A+A/438/769

J/A+AS/136/261 J/A+AS/136/531 J/A+AS/139/321 Tej A.: J/A+A/468/1001 J/MNRAS/381/1219 Tejfel V.G.: J/A+AS/139/47 Teker A.: J/A+A/394/187 J/A+A/447/685 Tektunali. G.H.: J/A+A/406/975 Tektunali H.G.: J/A+A/394/187 J/MNRAS/316/514 Telesco C.: J/ApJ/636/932 Televich R.J.: J/AJ/110/212 Telfer R.C.: J/ApJ/475/469 Telles E.: J/A+A/423/133 J/A+A/476/89 J/A+AS/129/455 J/AJ/110/212 Telleschi A.: J/A+A/468/353 J/A+A/468/379 TELLO C.: J/AJ/107/1381 TELTING J.H.: J/A+A/284/515 Telting J.H.: J/A+A/377/925 J/A+A/440/249 J/A+A/450/1149 Temirova A.V.: J/A+AS/101/447 J/A+AS/87/1 J/A+AS/91/337 J/AZh/73/835 J/AZh/79/589 J/other/BSAO/40.5

VIII/33 Temmer: J/A+A/390/707 Temmer M.: J/A+A/447/735 Tempirova A.V.: J/AZh/72/613 Templeton M.R.: J/AJ/130/776 J/other/IBVS/5526 Temporin S.: J/A+A/463/519 Ten Brummelaar T.: J/AJ/117/1890 ten Brummelaar T.A.: J/A+A/453/155 Ten Brummelaar T.A.: J/AJ/119/3084 Tennant A.F.: J/ApJ/563/L49 J/ApJ/615/887 J/ApJ/615/897 J/ApJS/144/213 J/ApJS/154/519 Tennyson J.: J/MNRAS/357/471 VI/119 VI/121 Tenorio-Tagle G.: J/A+AS/129/455 J/AJ/110/212 J/ApJ/619/270 Teodorani M.: J/A+A/367/199 Teodorescu A.M.: J/ApJ/635/290 Teoh A.: J/AJ/129/1993 Tepe A.: J/other/APh/26.282 Teplitz H.: J/AJ/112/2500 Teplitz H.I.: J/AJ/119/486 J/AJ/131/250 J/ApJ/548/127 J/ApJ/581/258 J/ApJ/614/75 J/ApJ/634/128

J/ApJ/634/L1 J/ApJS/142/79 J/ApJS/146/209 J/MNRAS/356/568 Teplyakova A.L.: J/A+A/442/109 ter Kuile C.R.: J/A+AS/128/179 Terashima Y.: J/ApJ/601/735

Terasranta H.: J/A+A/402/151 J/A+A/427/769 J/A+A/440/409 J/A+AS/116/157 J/A+AS/129/577

J/A+AS/132/305 Tercero B.: J/A+A/466/255 Terebbizh V.Y.: J/other/AISAO/32.31 Terebizh A.V.: J/A+AS/139/555 Terekhov O.V.: IX/36 J/A+AS/129/1 Tereshchenko V.M.: III/202 J/AZh/71/468 J/AZh/73/762 J/AZh/76/604 J/AZh/78/1135 J/AZh/79/249 Terlevich A.I.: J/MNRAS/330/547 Terlevich E.: J/AJ/110/212 J/ApJ/513/168 J/MNRAS/366/480 Terlevich R.: J/A+A/341/98 J/A+AS/91/285 J/MNRAS/366/480 Terlevich R.J.: J/ApJ/513/168 VII/176A Terndrup D.M.: J/AJ/109/1131 J/AJ/110/1774 J/AJ/110/2788 J/AJ/112/171 J/AJ/113/321 J/AJ/118/895

J/ApJ/353/494 J/ApJ/517/650 J/ApJS/143/73 TERNDRUP D.M.: J/MNRAS/272/391 Terra M.A.O.: J/A+AS/128/561 TERRANEGRA L.: J/A+A/299/89 Terranegra L.: J/A+A/341/L79 J/A+A/448/579 TERRANEGRA L.: J/A+AS/104/557 Terranegra L.: J/A+AS/123/329 J/A+AS/145/187 Terrell D.: J/AJ/110/2400 Terrier R.: J/ApJ/607/L33 J/ApJ/636/765 J/ApJS/170/175 Terry J.N.: J/A+A/383/398 J/A+A/393/57 Terzan A.: J/A+AS/120/275 J/A+AS/123/507 J/A+AS/49/715 J/A+AS/76/205 J/A+AS/90/451 Terzian Y.: J/AJ/127/501 J/ApJS/123/219 J/ApJS/89/189 Terzieva R.: J/ApJ/519/697 Testa V.: J/A+A/298/461 J/A+A/380/478 J/A+A/425/509 J/A+A/462/599 J/A+A/469/503 J/AJ/118/2839

J/AJ/121/916 TESTA V.: J/MNRAS/272/391

Testi L.: II/243 J/A+A/362/199 J/A+A/392/971 J/A+A/432/921 J/A+A/447/221 J/A+A/448/189

J/A+A/452/245 Testor G.: J/A+A/453/517 J/A+A/469/459 Testori J.C.: J/A+A/376/861 Teuben P.J.: J/ApJS/112/109 Teutsch P.: J/A+A/447/921 Thackeray A.D.: J/MNRAS/121/337 Thaddeus P.: J/A+A/315/578 J/A+A/357/1001 J/A+A/459/317 VIII/39 Thakar A.R.: J/AJ/121/2331 J/AJ/123/567 J/AJ/125/1817 J/AJ/126/2579 Thakkar D.D.: J/ApJS/98/33 Thanjavur K.: J/A+A/415/289 Thatte N.: J/A+A/443/41 Thatte N.A.: J/A+A/406/593 The DENIS Consortium: B/denis II/240 II/252 the ISO Team: VI/101 The Las Campanas Infrared Survey.: VII/228 The L.-S.: J/ApJ/611/452 the MACHO Team: II/247 The P.S.: J/A+A/327/145 J/A+A/380/609 J/A+AS/104/315 THE P.S.: J/A+AS/105/311 The P.S.: J/A+AS/119/1 J/A+AS/121/223 J/A+AS/123/63 J/AZh/77/420 Theile I. Voigt H.H.: III/15 III/15 III/76A III/76A Theissen A.: J/A+AS/143/409 Thejll P.: J/A+A/317/689 Theodorou A.: J/A+A/435/437 Theuns T.: J/ApJ/596/768 Theureau G.: J/A+A/383/398 J/A+A/430/373 J/A+A/465/71 J/A+AS/126/471 J/A+AS/130/333 VII/237

VII/238 THEVENIN F.: III/165 Thevenin F.: III/193 J/A+A/423/353 J/A+A/459/669 J/A+A/464/373 J/A+A/478/197 Thiele U.: J/A+AS/109/537 Thierbach M.: J/A+A/418/1 Thiering I.: J/ApJS/117/319 Thilker D.: J/A+A/406/829 Thilker D.A.: J/A+A/417/421 J/ApJ/633/871 Thim F.: J/ApJ/551/973 J/PASP/117/37 Thimm G.J.: J/A+A/285/785 THIMM G.J.: J/A+A/297/61 Thoburn C.: I/99 I/99 Tholen D.J.: J/A+AS/125/399 Thollander L.: J/other/APh/26.282 Thom C.: J/A+A/431/143 J/ApJS/161/147 Thomas D.: J/A+A/395/431 J/A+A/407/423 J/A+A/415/123 J/ApJ/621/673 Thomas H.-C.: J/A+A/335/467 J/A+A/347/47 J/A+A/350/805 Thomas J.: J/A+A/469/575 Thomas: J/A+A/395/753 J/A+A/395/753 THOMAS N.G.: I/112 Thomas N.G.: I/79 Thomas P.A.: J/AJ/106/848 THOMAS P.G.: J/ApJ/397/520 Thomas R.C.: J/A+A/437/789 Thomas R.J.: J/ApJ/543/1016 THOMAS R.J.: J/ApJS/91/461 Thomas S.: J/A+A/450/681 Thomas T.: J/A+A/446/19 Thomasson M.: J/AJ/121/182 Thome J.M.: I/114 THOME J.M.: I/88 Thommes E.: J/A+AS/120/267 Thompson A.B.: J/AJ/101/1408 Thompson D.: J/A+A/371/833 J/AJ/120/2331 J/ApJ/586/765 Thompson D.J.: IX/9 J/ApJS/101/259 J/ApJS/107/227 J/ApJS/123/79 J/ApJS/157/324 J/ApJS/94/551 Thompson G.I.: II/59B II/86 THOMPSON G.I.: III/39A V/38 Thompson I.: J/A+AS/119/413 J/AJ/124/1486 J/AJ/132/137 Thompson I.B.: J/A+A/373/899 J/A+A/424/1101 J/AJ/123/855 J/AJ/128/1606 J/AJ/132/85 Thompson M.A.: J/A+A/453/1003 J/A+AS/133/29 J/MNRAS/363/405 J/MNRAS/368/1223 Thompson R.I.: J/AJ/117/17 J/AJ/130/1 J/ApJ/540/1016 J/ApJ/546/694 J/ApJ/596/748 J/ApJ/606/L25

J/ApJ/624/L5 Thompson R.J.: J/AJ/115/2587 J/ApJ/434/54 Thompson S.A.: J/MNRAS/335/291 Thomson J.R.: J/A+A/394/97 J/AJ/130/767 Thorel J.C.: J/A+AS/115/59 Thorel J.-C.: J/A+AS/132/29 J/A+AS/147/251 Thoren P.: J/A+A/425/187 Thorne A.P.: J/ApJS/94/221 Thornley M.D.: J/ApJ/633/871 Thorsett S.E.: J/AJ/121/1192 Thorstensen J.: J/ApJ/470/724 Thorstensen J.R.: J/AJ/109/1458 J/AJ/109/2368 J/AJ/122/2893 J/AJ/126/3017 J/PASP/118/1238 Thorstensen R.: J/A+A/466/1031 Thorwirth S.: J/A+A/398/L11 J/A+A/425/767 Thoul A.: J/A+A/415/241 Thrall H.: J/A+A/472/805 Thronson H.A.: J/ApJ/452/238 Throop H.: J/ApJ/549/578 Throwirth S.: J/A+A/466/989 Thuan T.X.: J/A+A/421/539 J/A+A/448/955 J/A+AS/139/1 J/AJ/121/169 J/AJ/121/66 J/AJ/123/1292

J/AJ/127/1959 J/ApJ/241/486 J/ApJ/378/65 J/ApJ/511/639 J/ApJ/517/148 J/ApJ/602/200 J/ApJS/129/547 J/ApJS/161/240 J/ApJS/81/5

Thuet-Fleck E.: V/53A Thuillot W.: J/A+A/380/727 J/A+A/391/767 J/A+A/400/1095 J/A+A/406/751 J/A+AS/125/399 THUM C.: J/A+A/281/161 Thurman J.: IX/13 Thurmes P.M.: II/254 Thurston M.R.: J/A+A/375/863 Tian B.: J/A+A/455/247 TIAN K.P.: I/215 Tian K.P.: J/A+A/381/464 J/A+AS/100/243 J/A+AS/105/15 J/A+AS/118/503 Tian K.-P.: J/A+AS/128/255 Tian K.P.: J/A+AS/131/239 Tian K.-P.: J/A+AS/131/89 Tian K.P.: J/A+AS/133/387 Tian W.W.: J/A+AS/136/571 Tian W.-W.: J/other/AcApS/16.103

Tiede G.P.: J/AJ/109/1154 J/AJ/110/2788 J/AJ/113/321 J/AJ/118/895 J/AJ/127/213 J/AJ/128/224

J/ApJ/541/977 J/ApJS/143/73 Tielens A.G.G.M.: J/A+A/357/572 J/A+A/382/184 TIELENS A.G.G.M.: J/ApJS/86/713 Tielens: J/A+A/431/523 Tiersch H.: J/ApJ/523/87 VII/196 VII/89B VII/89B Tifft W.G.: VII/1B Tijani A.H.: J/A+A/411/565 Tikhomirova Y.: J/ApJ/563/80 Tikhonov A.V.: J/PAZh/30/243 Tikhonov N.: J/A+AS/142/347 Tikhonov N.A.: J/A+A/389/439 J/A+AS/119/499 Tilav S.: J/other/APh/26.282 Timothy J.G.: J/AJ/117/548 J/AJ/131/1000 Tingay S.J.: J/ApJS/155/33 J/ApJS/99/637 J/PASJ/52/997 J/PASJ/55/351 Tinker M.C.Jr: J/ApJ/632/L131 Tinney C.: J/ApJ/348/98 Tinney C.G.: J/A+A/338/1066 J/AJ/127/2856 J/AJ/130/1871 TINNEY C.G.: J/ApJ/414/279 Tinney C.G.: J/MNRAS/301/1031 J/MNRAS/370/163 J/MNRAS/372/163 Tinti S.: J/A+A/426/463 J/A+A/450/959 Tiphene D.: J/A+A/363/62 J/A+AS/135/133 Tirado H.: J/AJ/112/2408 Tisserand P.: J/A+A/454/185 Titov O.A.: I/266 J/AZh/81/1032 Titterington D.: J/MNRAS/341/1066 J/MNRAS/360/340 Titterington D.J.: J/MNRAS/246/110 Tjin A Djie H.R.E.: J/A+A/380/609 J/A+AS/124/33 J/MNRAS/310/210 Tkachenko A.Yu: IX/36 Tluczykont M.: J/A+A/390/39 J/A+A/406/L9 J/A+A/410/813 T.M.: J/A+A/386/1001 Toale P.A.: J/other/APh/26.282 Tobin W.: J/A+A/303/137 TOBIN W.: J/A+AS/107/385 J/A+AS/109/447 Tobin W.: J/MNRAS/304/733 J/MNRAS/331/609 J/MNRAS/349/833 Todd I.: J/MNRAS/362/1006 Tofani G.: J/A+A/368/845 J/A+AS/103/541 J/A+AS/134/115 Toft S.: J/AJ/130/1 J/ApJS/162/1 Tokarz S.: J/A+A/369/57 J/AJ/129/363 J/ApJ/510/659 Tokarz S.P.: J/PASP/111/438 TOKARZ S.P.: VII/144 Tokarz S.P.: VII/164 VII/193 Tokovinin A.: J/A+A/382/92 J/A+A/409/245 J/A+A/450/681 Tokovinin A.A: B/sb9 Tokovinin A.A.: III/191 J/A+A/282/831 J/A+A/314/846 J/A+A/374/227 J/A+A/382/118 J/A+A/384/1030

J/A+A/465/257 J/A+AS/121/71 J/A+AS/124/75 J/A+AS/136/373 J/PAZh/24/343 J/PAZh/24/868 J/PAZh/24/918

Tokovinin A.A: V/122 Tokunaga A.: J/ApJS/118/177 TOKUNAGA A.T.: J/AJ/106/656 Tokunaga A.T.: J/AJ/125/1537 J/ApJS/161/154 Tolchenov R.N.: VI/119 Tolea A.: J/AJ/125/1711 Toller G. N.: J/A+AS/127/1 Tollestrup E.V.: J/AJ/131/527 J/ApJS/96/117 Toloba E.: J/MNRAS/374/664 Tolstoy E.: J/A+A/459/423 J/AJ/116/146 J/AJ/123/3154 J/AJ/125/1261 J/AJ/127/2723 J/AJ/128/1177

J/AJ/129/1465 J/MNRAS/327/918 Tomaney A.: J/A+A/345/521 J/ApJ/542/257 Tomaney A.B.: J/ApJ/552/289 Tomasella L.: J/A+A/343/806 J/A+A/415/145 J/A+AS/137/521 Tomasi P.: J/A+AS/43/1 Tomassetti G.: VIII/37 Tomczyk S.: J/A+A/390/717 Tomei M.: J/A+A/469/1211 Tomita A.: J/AJ/111/42 J/ApJS/89/57 Tomita H.: J/ApJS/141/23 J/ApJS/141/31 Tomkin J.: J/A+A/275/101 J/A+A/327/587 J/A+AS/102/603 J/MNRAS/340/304 Tomov T.: J/A+A/434/1107 J/A+AS/116/1 J/A+AS/122/43 Tomov T.V.: J/A+A/311/484 Tompkins W.: J/ApJS/107/227 Tompkins W.F.: J/ApJS/123/79 Tonello N.: J/A+A/410/813 Toner R.: J/ApJS/162/1 Tonry J.: J/ApJ/607/665 VII/185 Tonry J.L.: J/AJ/108/2087 J/AJ/115/1016 J/AJ/115/1894 J/AJ/117/2810 J/AJ/118/346 J/AJ/130/2453

J/AJ/131/1648 J/ApJ/594/1 J/ApJ/602/571 J/ApJ/613/200 J/ApJ/613/279 J/ApJS/153/223 J/ApJS/154/509 J/MNRAS/327/1004 VII/72

Toque N.: J/A+A/475/1053 Torii K.: J/other/GCN/1070.1 TORISEVA M.: J/A+AS/111/95 Toriseva M.: J/A+AS/133/337 Tormen G.: J/ApJS/96/123 Torniainen I.: J/A+A/469/451 J/AJ/133/1947

Tornikoski M.: J/A+A/427/769 J/A+A/445/441 J/A+A/469/451 J/A+AS/116/157 J/A+AS/129/577

J/A+AS/132/305 J/AJ/133/1947 Torppa J.: J/A+A/350/513 J/A+A/427/769 Torra J.: J/A+A/319/881 J/A+A/326/1044 J/A+A/326/722 J/A+A/336/953 J/A+A/375/539 J/A+A/425/489

J/A+A/446/785 J/A+AS/115/401 J/A+AS/128/265 Torrejon J.M.: J/A+AS/119/271 J/A+AS/123/1 Torrelles J.M.: J/A+A/470/231 J/A+A/471/573 J/ApJ/628/789 Torrence G.W.: VIII/42 Torres A.V.: III/136

Torres C.: J/A+AS/105/179 J/A+AS/110/27 J/A+AS/115/481 J/A+AS/122/447 J/A+AS/128/457

J/A+AS/78/231 J/A+AS/83/317 J/A+AS/91/129 J/A+AS/92/449 Torres C.A.O.: J/A+A/460/695 J/A+AS/128/561 J/AJ/126/2971 Torres D.F.: J/A+A/370/468 Torres G.: B/sb9 J/A+A/325/647 J/A+A/375/130 J/A+A/375/L27 J/A+A/456/651 J/AJ/119/1389

J/AJ/124/1144 J/AJ/129/2294 J/AJ/129/2798 J/AJ/130/234 J/AJ/130/2838 J/AJ/131/1702 J/AJ/131/2274 J/ApJ/640/1018 V/122

Torres-Dodgen A.V.: III/143 III/181 J/ApJ/446/300 Torres-Peimbert S.: III/84B TORRES-PEIMBERT S.: J/AJ/105/35 Torres-Peimbert S.: J/ApJ/513/168 J/ApJS/120/113 TOSI M.: J/AJ/105/894 Tosi M.: J/AJ/109/173 J/AJ/121/327 J/MNRAS/299/834 J/MNRAS/309/739 Tosti G.: J/A+A/327/61 J/A+A/395/11 J/A+A/402/151 J/A+A/423/935 J/A+A/469/503 J/A+AS/116/403

J/A+AS/129/577 J/A+AS/130/109 Toth L.V.: J/A+A/418/131 J/A+A/453/923 J/A+A/463/1227 Totten E.J.: J/MNRAS/335/291 J/MNRAS/351/1401 Tougou N.: J/PASJ/52/631 Tout C.A.: J/AJ/128/1177 J/MNRAS/319/215 Tovmassian H.M.: J/AJ/124/1283 J/ApJ/523/87 Townsley L.: J/AJ/120/1426 J/ApJ/574/258 J/ApJ/575/354 J/ApJ/584/911 J/ApJS/160/319 J/ApJS/160/353

J/ApJS/163/306 Townsley L.K.: J/AJ/122/2810 J/AJ/126/539 J/AJ/131/2164 J/ApJ/554/742 J/ApJ/589/225 Toyoda S.: J/PASJ/53/1081 Tozzi P.: J/A+A/432/381 J/A+A/451/457 J/AJ/123/2223 J/AJ/127/213 J/ApJ/595/685 J/ApJ/633/174

J/ApJS/139/369 J/ApJS/155/271 J/ApJS/155/73 J/ApJS/161/21 T.P.: J/A+A/364/391 T.Pursimo: J/A+A/427/769 Tracewell D.A.: V/102 V/105 V/109 V/95 Trachternach C.: J/A+A/458/341 Trafton L.M.: J/AJ/116/941 Trager S.C.: J/A+A/453/493 J/AJ/109/218 J/ApJS/116/1 Trams N.: J/A+A/365/545 Trams N.R.: J/A+A/307/869 J/A+A/310/893 Tran H.D.: J/AJ/127/3121 J/AJ/128/1990 J/ApJ/612/202 J/ApJ/621/53 J/ApJS/150/1 Tran K.-V.: J/ApJS/140/143 Tran K.-V.H.: J/ApJ/549/172 J/ApJ/644/30 J/ApJ/661/750 TRAN MINH F.: J/A+A/297/494 Trancho G.: J/A+A/435/65 Trasarti-Battistoni R.: J/A+AS/130/341 Traub W.A.: J/ApJ/620/961 Trauger J.: J/AJ/113/1001 Trauger J. T.: J/PASP/108/682 Trauger J.T.: J/AJ/117/1700 TREFFERS R.R.: J/AJ/107/1022 Treister E.: J/AJ/131/2373 J/AJ/134/1103 J/ApJ/621/104 J/ApJ/640/603 J/ApJS/162/1 Tremaine S.: J/AJ/129/2138 TREMKO J.: J/A+A/285/459 Tremko J.: V/119 Tremonti C.: J/AJ/121/2308 J/AJ/121/2331 J/PASJ/55/771 Trentham N.: J/AJ/130/1502 J/MNRAS/333/423 J/MNRAS/335/712 J/MNRAS/369/1375 Trentini S.: J/A+A/321/123 Tresse L.: J/A+A/403/857 J/A+A/428/1043 J/A+A/457/79 J/A+A/467/73 J/ApJ/481/49 J/ApJ/499/112

VII/225A Treu T.: J/ApJ/591/53 J/ApJ/627/32 J/ApJ/633/174 J/ApJ/634/977 Treves A.: J/A+A/327/61 J/A+A/469/503 Trevese D.: J/A+A/469/1211 J/A+AS/110/313 J/A+AS/125/459 J/A+AS/146/373 J/A+AS/94/327 Treyer M.A.: J/MNRAS/300/303 J/MNRAS/312/442 J/MNRAS/330/75 TRIAY R.: VII/32 Trigilio C.: J/A+A/379/755 J/PASJ/52/997 Trigueiros A.G.: J/ApJS/121/591 Trilling D.: J/ApJ/620/1010 Trilling D.E.: J/ApJ/636/1098 Trinchieri G.: J/A+A/330/423 J/A+A/426/11 J/A+A/435/799 J/A+A/439/413 Trinchieri G: J/A+A/448/1247 Trinchieri G.: J/A+A/467/73 J/A+AS/137/101 J/AJ/129/1832 J/ApJ/634/272 J/ApJS/80/531 J/MNRAS/382/279 Tripathi D.: J/A+A/422/337 Tripp T.M.: J/ApJ/508/200 J/ApJ/636/631 J/ApJS/155/351 Trippe S.: J/ApJ/643/1011 Trisoglio M.: J/A+A/374/1049 Tritton S.B.: J/MNRAS/362/689 VII/191 Troja E.: J/A+A/474/793 Troland T.H.: J/ApJS/145/329 J/ApJS/151/271 Troller M.: J/AJ/133/1947 Trombka J.: J/ApJS/156/217 Trosell E.: J/ApJS/162/428 TROUP E.: J/MNRAS/251/330 Troup E.: VIII/38 Troup E.R.: J/MNRAS/301/1019 TRUDEL J.-L.: J/AJ/105/2291

Truemper J.: II/208 IX/10A J/A+A/307/33 J/A+A/317/328 J/A+A/323/739 J/A+A/329/482 J/A+A/335/467 J/A+A/347/47 J/A+A/371/833 J/A+A/373/63 J/A+AS/106/303 J/A+AS/109/147 J/A+AS/114/215

Truen-Liang Lee J.: I/174 Trujillo I.: J/AJ/122/38 J/ApJ/602/664 J/ApJ/604/521 J/MNRAS/355/82 Trullols E.: J/A+A/324/549 J/A+A/333/471 J/A+A/337/125 TRULLOLS E.: J/A+AS/114/489 Trullols E.: J/A+AS/131/239 J/A+AS/146/169 TRULLOLS E.: J/A+AS/94/519 Trump J.R.: J/AJ/126/2579 J/ApJS/165/1 Trumper J.: IX/29 IX/32 J/A+A/278/36 J/AJ/126/2209 J/ApJ/450/51 J/ApJS/129/435

Trundle C.: J/A+A/419/1123 J/A+A/437/467 J/A+A/456/623 J/A+A/457/265 J/A+A/466/277

J/A+A/471/625 Trundle: J/A+A/434/1125 Trushkin S.A.: J/AZh/78/922 J/other/BSAO/46.62 J/other/BSAO/50.115 Tsalmantza P.: J/A+A/447/89

Tsapras Y.: J/A+A/365/110 J/A+A/378/116 J/A+A/379/564 J/MNRAS/343/1131 J/MNRAS/351/1071

J/MNRAS/358/795 J/MNRAS/362/117 Tsay W-S.: J/AJ/112/628 Tsay W.-S.: J/AJ/117/2757 Tschager W.: J/A+A/402/171 J/A+A/435/863 Tsoti G.: J/A+AS/117/475 Tsuboi Y.: J/AJ/120/1426 J/AJ/125/1537 J/AJ/130/2212 J/ApJ/566/974 J/ApJ/574/258 J/ApJ/605/259

J/ApJ/626/272 J/PASJ/55/653 Tsuchiyama A.: J/A+A/449/583 Tsuji T.: J/A+A/310/933 J/A+A/328/175 J/A+A/353/528 Tsujimoto M.: J/AJ/125/1537 J/AJ/130/2212 J/ApJ/566/974 J/ApJ/635/214 J/ApJS/160/319 J/ApJS/163/306

J/PASJ/55/161 J/PASJ/55/635 J/PASJ/55/653 Tsuru T.: J/ApJ/518/656 Tsuru T.G.: J/PASJ/57/135 Tsuruta S.: J/PASJ/52/631 Tsuzuki Y.: J/A+A/365/409 Tsvetanov Z.: J/AJ/122/1330 J/AJ/123/3409 J/AJ/125/1711 J/ApJ/600/L99 J/ApJ/607/665 Tsvetanov Z.I.: J/AJ/127/3553 J/AJ/128/1990 J/ApJ/483/745 J/ApJ/612/202 J/ApJ/621/53 J/ApJS/101/287

J/ApJS/150/1 TSVETKOV D.Yu.: II/176 Tsvetkov D.Yu.: II/201 II/218 II/256 J/PAZh/30/37 Tsvetkov M.K.: VI/90 Tsvetkova K.P.: VI/90 Tsvetkova T.M.: II/172B II/205 J/PAZh/20/693 J/PAZh/21/509 J/PAZh/28/201 Tsvetkova V.S.: J/A+A/420/447 Tsymbal V.: J/A+A/457/1033 J/PASP/118/636 Tu C.: J/A+AS/126/503 Tubbesing S.: J/A+A/389/931 Tuccari G.: J/PASJ/52/997 Tucholke H.-J: I/221 Tucholke H.-J.: J/A+A/312/74 TUCHOLKE H.-J.: J/A+AS/104/161 J/A+AS/112/367 TUCHOLKE H.J.: J/A+AS/93/293 J/A+AS/93/311 Tucker D. (5): J/A+A/459/L1 Tucker D.: J/A+A/462/811 J/AJ/112/2408 Tucker D.L.: J/AJ/119/2598 J/AJ/123/1807 J/AJ/127/1811 J/AJ/127/1883 J/AJ/127/2210 J/AJ/129/2062

J/AJ/130/873 J/AN/321/101 J/AN/327/365 J/ApJS/130/237 VII/203 Tucker G.S.: J/ApJS/148/97 Tucker R.H.: I/144 Tueller J.: J/ApJ/633/L77 Tuemer T.: IX/9 Tuffs R.: J/A+AS/120/489 J/A+AS/121/507 J/A+AS/125/293 Tuffs R.J.: J/ApJ/619/755 Tufte S.L.: II/249 Tufts J.: J/MNRAS/344/283 Tufts J.R.: J/MNRAS/357/1231 Tully J.A.: J/A+A/365/266 J/A+A/430/331 Tully R.B.: J/AJ/112/2471 J/AJ/130/1502 J/ApJ/533/744 J/ApJ/580/774 J/ApJS/80/479 J/MNRAS/335/712

J/MNRAS/369/1375 Tully R.B: VII/145 Tully R.B.: VII/75 Tumer E.: J/ApJS/127/79 Tumer O.: J/A+AS/129/9 TUNCA Z.: J/A+A/281/811 J/A+A/287/817 Tunca Z.: J/A+A/336/920 J/A+AS/129/9 Tuncel S.: J/AN/326/321 TUOMINEN I.: J/A+A/278/449 J/A+A/282/L9

Tuominen I.: J/A+A/309/655 J/A+A/334/863 J/A+A/341/553 J/A+A/350/513 J/A+A/351/212 J/A+A/360/499 J/A+A/374/1049 J/A+A/380/578 J/A+A/390/179 J/A+A/394/505 J/A+A/402/745 J/A+A/432/657 J/A+A/476/881

TUOMINEN I.: J/A+AS/103/183 Tuominen I.: J/A+AS/123/209 J/A+AS/139/513 TURATTO M.: II/103 Turatto M.: II/159 II/227 TURATTO M.: J/A+A/297/49 Turatto M.: J/A+A/317/423 J/A+AS/65/349 J/MNRAS/328/1181 J/MNRAS/369/1880 Turcan D.: J/other/APh/26.282 Turck-Chieze S.: J/A+A/478/197 Turcu V.: J/A+AS/125/399 Turner A.: J/ApJ/463/26 J/ApJ/508/491 J/ApJ/521/155 J/ApJ/523/540 Turner A.M.: J/ApJ/512/48 TURNER B.E.: VI/56 Turner D.G.: J/AJ/113/2104 J/AJ/130/1194 J/ApJS/137/209 Turner E.L.: J/AJ/116/2081 Turner G.W.: J/ApJS/115/185 Turner M.J.L.: VII/19 Turner N.H.: J/A+A/453/155 J/AJ/108/2299 J/AJ/111/936 J/AJ/119/3084 J/ApJ/525/420 Turner T.J.: J/ApJ/582/615 Turnshek D.A.: J/ApJ/636/610 TURON C. et al.: I/191 Turon C.: I/196 I/197A J/A+A/331/81 J/A+A/345/471 J/A+A/430/165 Turriziani S.: J/A+A/472/699 Tustin A.W.: J/AJ/122/1289 TUTUKOV A.: V/41 Tutukov A.V.: J/A+A/288/713 J/AZh/79/738 Tuvikene T.: J/A+A/463/1053 TUZI K.: I/38 Twarog B. A.: J/AJ/119/2282 TWAROG B.A.: J/AJ/107/1577 Twarog B.A.: J/AJ/109/2828 J/AJ/116/1922 J/AJ/125/1383 J/AJ/127/1000 J/AJ/129/872 J/AJ/131/461

J/PASP/106/281 TWAROG B.A.: J/PASP/106/486 Twarog B.A.: J/PASP/107/32 Tylenda R.: J/A+A/318/256 J/A+A/327/736 J/A+A/373/1032 J/A+A/405/627 J/A+AS/106/559 V/100

V/84 Tyler K.D.: J/MNRAS/380/986 Tyson J.A.: J/AJ/116/102 Tyson N.: J/AJ/112/2408 Tytler D.: J/A+A/458/427 TYTLER D.: J/A+AS/108/287 J/AJ/106/426 Tytler D.: J/AJ/119/2571 Tyul’bashev S.A.: J/AZh/74/812 VIII/67 Tyurina N.V.: J/PAZh/27/250 Tyutyunnyk A.A.: J/other/KFNT/22.187 Tyynela J.: J/A+A/350/513 Tzioumis A.: J/AJ/115/1253 J/MNRAS/323/757

Tzioumis A.K.: J/AJ/127/3609 J/AJ/130/2529 J/MNRAS/298/692 J/MNRAS/325/1411 J/PASJ/52/997 J/PASJ/55/351

Ubertini P.: J/ApJ/607/L33 J/ApJ/636/765 J/ApJS/170/175 Uchida M.M.M.: J/A+A/423/199 Udalski A.: I/244A II/213 II/278 J/A+AS/112/407 J/A+AS/120/139 J/A+AS/122/471 J/A+AS/125/343

J/A+AS/128/19 J/AcA/48/113 J/AcA/48/147 J/AcA/48/175 J/AcA/48/563 J/AcA/49/223 J/AcA/49/319 J/AcA/49/437 J/AcA/49/521 J/AcA/49/543 J/AcA/50/1 J/AcA/50/337 J/AcA/50/355 J/AcA/51/175 J/AcA/51/221 J/AcA/51/259 J/AcA/51/317 J/AcA/52/129 J/AcA/52/143 J/AcA/52/217 J/AcA/53/1 J/AcA/53/93 J/AcA/54/347 J/AcA/56/253 J/AJ/110/1141 J/AJ/128/1167 J/AJ/128/2815 J/ApJ/636/240 J/MNRAS/348/1439 J/MNRAS/356/331 J/PASP/114/298

Udaltsov V.A.: J/AZh/71/110 Udry: J/A+A/404/775

Udry S.: B/sb9 J/A+A/356/590 J/A+A/364/665 J/A+A/375/205 J/A+A/375/L27 J/A+A/379/999 J/A+A/388/632 J/A+A/390/267 J/A+A/392/215 J/A+A/394/151 J/A+A/395/97 J/A+A/398/363 J/A+A/406/373 J/A+A/410/1039 J/A+A/410/1051 J/A+A/414/351 J/A+A/415/391 J/A+A/426/695 J/A+A/427/313 J/A+A/430/165 J/A+A/450/681 J/A+A/458/997 J/A+A/466/743 J/A+A/471/L51 J/A+A/474/293 J/A+A/475/1125 J/A+A/476/1347 J/A+AS/130/223 J/AJ/114/1087 V/117 V/122

UEda Y.: J/A+A/469/27 J/ApJ/518/656 J/ApJ/610/128 J/ApJS/133/1 J/ApJS/148/275 J/ApJS/161/185

J/MNRAS/357/1357 J/MNRAS/372/741 J/PASJ/53/445 UEhara H.: VII/244A UEmura M.: J/A+A/463/1053 UEno M.: J/ApJ/635/214 UEsugi A.: III/63B UEta T.: J/ApJS/122/221 Ugarte P.: J/AJ/112/2408 Uglova I.M.: J/PAZh/23/674 J/PAZh/23/90 Ugolkova L.S.: J/other/BSAO/45.136 Ugryumov A.: J/A+A/419/469

Ugryumov A.V.: J/A+A/366/771 J/A+A/373/24 J/A+A/374/907 J/A+A/397/463 J/A+A/442/109

J/A+AS/131/295 J/A+AS/135/511 J/A+AS/137/299 J/A+AS/142/417 Ulaczyk K.: J/A+A/440/783 Ulla A.: J/A+A/458/173 Ullan A.: J/A+A/476/1191 Ulmer M.: J/A+A/451/1159 J/A+A/459/679 Ulmer M.P.: J/AJ/120/1 J/ApJS/126/209 ULMER M.P.: J/ApJS/56/403 Ulrich D.: J/AJ/112/1188 Ulvestad J.S.: J/ApJ/609/564 J/ApJS/133/77 Umana G.: J/A+A/379/755 Umeda K.: J/PASJ/57/881 Umemoto T.: J/PASJ/52/997 VII/244A Umemura M.: J/ApJ/648/7 J/PASJ/57/165 Ungerechts H.: J/A+A/402/151 J/A+AS/122/271 VIII/39 Ungruhe R.: J/A+AS/125/71 J/A+AS/93/463 Uomoto A.: J/AJ/121/2308 J/AJ/121/2331 J/AJ/123/567 J/AJ/124/2164 J/AJ/126/1023 J/AJ/126/2579

J/AJ/127/3553 J/ApJ/607/426 Upadhyay K.: J/MNRAS/358/1290 UPGREN A.R.: III/10A III/11 III/12 III/9 Upgren A.R.: J/AJ/77/486 Urban S: I/259 Urban S.: I/270

Urban S.E.: I/171 I/219 I/222 I/223 I/224 I/225 I/226 I/227 I/228 I/231 I/232 I/246 I/247 I/268 I/275 I/276 I/289 I/294 I/297 J/AJ/120/1120 J/AJ/122/1586 J/AJ/124/2254 J/AJ/127/539 J/AJ/128/3012 J/AJ/132/50

Urbaneja M.A.: J/A+A/422/511 J/A+A/437/467 Uribe A.: J/PASP/114/233 Urick V.J.: J/ApJS/140/331 Urosevic D.: J/A+A/435/437 Urpo S.: J/A+AS/132/305 Urquhart J.S.: J/A+A/461/11 J/A+A/474/891 Urry C.M.: J/AJ/131/2373 J/AJ/134/1103 J/ApJ/621/104 J/ApJ/640/603 J/ApJS/162/1 Usenko I.A.: J/A+A/281/465 J/A+A/305/551 Usher P.D.: II/115A II/81 J/AJ/111/645 J/ApJS/153/119 USHER P.D.: VII/63 Uylings P.H.M.: J/A+A/340/300 J/A+AS/125/539 J/A+AS/130/541

Uytterhoeven K.: J/A+A/415/241 J/A+A/438/201 J/A+A/440/249 J/A+A/446/583 J/A+A/449/305

J/A+A/463/243 Uzpen B.: J/ApJ/635/560

V.: J/A+A/428/587 J/A+A/446/949 J/A+A/469/529 J/AZh/79/328 V. K.: J/AZh/72/291 Vacca W.D.: J/A+A/406/593 J/A+A/440/61 J/AJ/113/1691 J/AJ/117/1708 Vaccari M.: J/AJ/129/2074 J/MNRAS/351/1290 J/MNRAS/355/97 J/MNRAS/358/333 J/MNRAS/358/397

J/MNRAS/372/1621 Vaeisaenen P.: J/A+A/463/1029 Vagnetti F.: J/A+A/469/1211 Vaiana G.S.: J/ApJ/315/687 J/ApJ/325/798 J/ApJ/348/253 J/ApJ/348/557 J/ApJ/351/492 Vaillancourt J.E.: J/ApJ/604/717 Vaisaenen P.: J/A+AS/112/99 J/A+AS/119/153 Vaisanen P.: J/A+A/474/473 J/AJ/131/527 J/MNRAS/351/1290 J/MNRAS/358/397 Vakili F.: J/A+A/312/879 Vakulik V.G.: J/A+A/420/447 Val Baker A.: J/A+A/441/685 Valbousquet A.: I/2 VALBOUSQUET A.: I/21A I/21A I/22A I/22A Valcheva A.T.: J/A+A/466/501 Valdes F.: J/A+A/364/102 J/ApJS/152/251 Valdes J.R.: J/A+A/376/775 J/A+A/422/925 J/A+A/434/149 J/AJ/124/1283 Valdettaro R.: J/A+A/327/755 J/A+A/368/845 Valdez-Gutierrez M.: J/AJ/124/3157 Valenti E.: J/A+A/419/139 J/A+A/455/271 J/A+A/464/953 J/MNRAS/351/1204 J/MNRAS/361/272 Valenti J.: J/A+A/415/331 J/ApJ/622/1102 Valenti J.A.: J/ApJS/147/305 J/ApJS/159/141 VALENTIJN E.A.: J/A+AS/108/621 Valentijn E.A.: VII/115 Valentinuzzi T.: J/A+AS/116/1 Valiron P.: J/A+A/472/1029 VALITOVA A.M.: I/150 Valladares G.: J/AJ/112/2408 Vallbe M.: J/MNRAS/349/135 Vallee J.P.: J/A+AS/77/31 VALLENARI A.: J/A+A/300/128 Vallenari A.: J/A+A/372/879 J/A+A/434/1107 J/A+A/452/179 VALLENARI A.: J/A+AS/108/571 J/A+AS/112/367 Vallerga J.V.: J/AJ/107/751 J/AJ/108/1843

Valtaoja E.: J/A+A/445/441 J/A+AS/116/157 J/A+AS/129/577 J/A+AS/132/305 J/A+AS/143/357

J/AJ/133/1947 Valtchanov I.: J/A+A/363/440 J/A+A/392/795 J/MNRAS/382/279 VII/190 VII/201 Valtier J.C.: J/A+A/304/406 Valtier J.-C.: J/A+A/448/697 Valtonen: J/A+AS/128/459 Valtonen M.: J/A+AS/132/305 VALTONEN M.J.: J/ApJ/413/453 Valtonen T.: J/A+A/427/769 J/A+AS/132/305 Val’tts I: J/AZh/79/328 Val’tts I.E.: J/A+AS/124/85 J/A+AS/134/115 Valtts I.E.: J/AZh/71/37 J/AZh/72/22 van Albada G.B.: II/221A Van Albada T.S.: J/A+A/442/137 J/A+AS/115/407 van Albada T.S.: J/A+AS/116/15 VAN ALBADA-VAN DIEN E.: I/128 Van Altena W.: J/AJ/127/1158 Van Altena W.F.: I/174 VAN ALTENA W.F.: I/212 van Altena W.F.: I/238A I/277 I/288 Van Altena W.F.: J/AJ/117/548 van Altena W.F.: J/AJ/122/1486 J/AJ/124/601 Van Altena W.F.: J/AJ/126/2922 van Altena W.F.: J/AJ/126/762 Van Altena W.F.: J/AJ/127/1727 J/AJ/131/1000 van Belle G.T.: J/PASP/114/656 van Breugel W.: J/A+A/362/519 J/A+AS/143/303 J/AJ/123/1784 J/AJ/123/637 Van Breugel W.: J/ApJ/505/L7 J/ApJS/107/19 J/ApJS/131/185 van Breugel W.J.M.: J/A+A/431/793 Van Breugel W.J.M.: J/A+A/461/823 J/AJ/126/2237 Van Breukelen C.: J/MNRAS/379/1343 Van Buren D.: J/PASP/105/1127 Van Cauteren P.: II/242 J/A+A/374/1056 J/A+A/418/249 J/A+A/438/201 VAN DE GRIFT E.: J/A+AS/85/915 van de Steene G. C.: J/A+A/373/536 Van De Steene G.: J/A+A/419/563 VAN DE STEENE G.C.: J/A+A/299/238 J/A+AS/108/485 Van de Steene G.C.M.: J/A+A/274/895 van de Ven G.: J/A+A/445/503 Van De Wel A.: J/AJ/125/1107 J/ApJ/604/521 Van den Ancker M.E.: J/A+A/352/228 van den Ancker M.E.: J/A+A/380/609 J/A+AS/119/1 J/A+AS/121/223 J/A+AS/123/63 J/A+AS/124/33

J/MNRAS/310/210 van den Berg M.: J/A+A/382/899 Van Den Berg M.: J/A+A/418/509 VAN DEN BERGH S.: II/133 van den Bergh S.: J/AJ/112/359 Van den Bergh S.: J/AJ/117/247 van den Bergh S.: J/AJ/120/2190 J/AJ/122/611 J/AJ/122/621 J/AJ/124/782 Van Den Bergh S.: J/AJ/130/1194 van den Bergh S.: J/AJ/131/1559 J/AJ/84/1305 J/ApJ/588/218 Van Den Bergh S.: J/ApJ/590/797 J/ApJ/627/621 van den Bergh S.: J/ApJS/107/1 J/ApJS/92/219 Van Den Bergh S.: J/MNRAS/274/1107 van den Bergh S.: J/PASP/114/820 J/PASP/115/1280 J/PASP/117/773 VII/21 VII/218 VII/87A van den Besselaar E.J.M.: J/A+A/466/1031 van den Bosch F.C.: J/AJ/108/1598 van den Bosch R.C.E.: J/A+A/445/503 van den Heuvel E.P.J.: J/A+A/331/949 J/A+A/442/1135 J/A+A/455/1165 J/A+A/469/807 J/A+A/473/523

V/106 VAN DEN HOEK L.B.: J/A+AS/104/179 van den Hoek L.B.: J/A+AS/123/305 van der Bliek N.S.: J/A+A/453/101 Van der Bliek N.S.: J/A+AS/119/547 VAN DER HUCHT K.A.: III/141 van der Hucht K.A.: III/215 III/85 J/A+A/331/949 Van der Hucht K.A.: J/A+A/384/1012 Van der Hucht K.A: J/A+A/404/L29 van der Hucht K.A.: J/A+A/477/593 J/A+AS/119/37 J/ApJS/149/123 Van Der Hulst J.M.: J/A+A/442/137 Van der Hulst J.M.: J/A+A/453/493 VAN DER KRUIT P.C.: J/A+AS/106/451 van der Laan H.: J/A+A/374/861 van der Marel R.P.: J/AJ/124/3255 Van Der Marel R.P.: J/AJ/125/478 J/AJ/127/105 J/ApJS/161/304 van der Meer A.: J/A+A/473/523 van der Tak: J/A+A/432/369 VAN DER VEEN W.E.C.J.: J/A+AS/93/121 van der Walt D.J.: J/A+AS/110/81 van der Werf P.: J/AJ/122/2205 Van Der Werf P.: J/AJ/125/1107 J/AJ/131/1891 J/ApJ/604/521 J/MNRAS/351/1290 Van Der Wolf F.L.A.: J/A+A/306/924 van der Wolf F.L.A.: J/A+A/310/893 VAN DESSEL A.: J/A+AS/100/517 van Dessel E.: J/A+A/402/1157 J/A+AS/124/353 J/A+AS/136/525 J/A+AS/139/69 van Dessel E.L.: J/A+AS/112/291 J/A+AS/119/483 van det Tak F.F.S.: J/A+A/475/549 van Dijk R.: J/A+AS/143/145 Van Dishoeck E.F.: J/A+A/279/541 van Dishoeck E.F.: J/A+A/323/469 J/A+A/432/369 J/A+A/475/549 J/A+AS/124/205 van Dokkum P.: J/AJ/122/2205 Van Dokkum P.: J/AJ/130/1 J/AJ/134/1103 J/ApJ/498/195 J/ApJ/539/577 J/ApJ/606/L25 J/ApJ/661/750 Van Dokkum P.G.: J/AJ/125/1107 van Dokkum P.G.: J/AJ/130/2647

Van Dokkum P.G.: J/AJ/131/1891 J/ApJ/500/714 J/ApJ/604/521 J/ApJ/608/L93 J/ApJ/633/174

J/ApJS/162/1 van Driel W.: J/A+A/365/1 J/A+A/386/140 J/A+A/405/99 J/A+A/408/465 J/A+A/408/67 J/A+A/428/823 Van Driel W.: J/A+A/449/929 van Driel W.: J/AJ/130/2613 Van Driel W.: J/MNRAS/352/478 J/MNRAS/357/819 van Duinen R.J.: II/60 II/97 Van Duyne J.: J/AJ/127/1959 J/ApJ/640/603 Van Dyk S.: II/246 Van Dyk S.D.: J/AJ/116/2341 van Dyk S.D.: J/AJ/122/1844 van Dyke Dixon W.: J/AJ/112/1335 Van Eijndhoven T.J.: J/other/APh/26.282 van Flandern T.C.: I/163 van Genderen A.: J/A+A/384/1012 J/A+A/404/L29 van Genderen A.M.: J/A+A/304/415 Van Genderen A.M.: J/A+A/343/847 van Genderen A.M.: J/A+A/366/508 J/A+A/397/921 J/A+AS/119/37 J/A+AS/98/505 Van Haarlem M.P.: J/MNRAS/339/913 VAN HAMME W.: J/AJ/106/2096 Van Hamme W.: J/AJ/110/1350 J/AJ/128/1319 J/AJ/129/2798 Van Hamme W.V.: J/ApJ/484/394 van Heerde G.M.: J/A+AS/58/1 van Hoof P.A.M.: J/A+A/416/383 J/A+A/424/363 J/A+A/433/745 J/ApJS/149/157 van Houten C.J.: J/A+A/369/527 J/A+AS/109/425 J/A+AS/137/113 Van Kampen E.: J/MNRAS/372/1621 Van Kerckhoven C.: J/A+A/463/243 VAN KERKWIJK M.H.: J/A+A/290/609 van Kerkwijk M.H.: J/A+A/377/925 J/A+A/416/1037 Van Kerkwijk M.H.: J/A+A/439/433 van Kerkwijk M.H.: J/A+A/473/523 J/ApJ/563/L49 van Langevelde H.J.: J/A+A/366/481 J/A+A/391/967 J/A+A/472/805 J/A+AS/128/35 van Leeuwen F.: I/207 J/A+A/360/472 J/A+AS/105/15 J/A+AS/118/503 J/A+AS/130/157 Van Leeuwen F.: J/AJ/127/980 van Leeuwen J.: J/ApJS/107/239 van Loon J.: II/202 J/A+A/362/215 van Loon J.T.: J/A+A/406/51 J/A+AS/119/1 Van Loon J.T.: J/ApJ/552/289 Van Loon J.TH.: J/MNRAS/376/313 Van Malderen R.: J/A+A/463/243 van Paradijs J.: II/208 J/A+A/242/401 J/A+A/265/77 Van Paradijs J.: J/A+A/278/36 J/A+A/312/93 van Paradijs J.: J/A+A/317/328 Van Paradijs J.: J/A+A/325/1115

van Paradijs J.: J/A+A/373/63 J/A+A/377/925 J/A+A/442/1135 J/A+A/455/1165 J/A+A/469/807

J/A+AS/101/87 J/A+AS/114/215 J/A+AS/118/545 J/A+AS/126/401 Van Paradijs J.: J/A+AS/128/29 J/ApJS/134/385 van Paradijs J.: V/106 V/90 van Raai M.: J/A+A/466/641 van Starkenburg L.: J/AJ/122/2205 Van Starkenburg L.: J/AJ/125/1107 J/ApJ/604/521 VAN VEERDEN J.E DE VRIES J.: II/148 van Winckel H.: III/174 J/A+A/307/869 J/A+A/310/164 J/A+A/336/263 Van Winckel H.: J/A+A/417/269 J/A+A/449/305 J/A+A/463/243 van Winkler K.: VI/34 van Woerden H.: VII/75 Van Wyk F.: J/AJ/132/111 van Wyk F.: J/MNRAS/334/20 J/other/SAAOC/15.65 J/other/SAAOC/15.78 Van Zee L.: J/AJ/116/2805 van Zee L.: J/AJ/130/2613 J/ApJ/636/214 Vanautgaerden J.: J/A+A/463/243 Vanbeveren D.: J/A+AS/103/67 Vance S.J.: J/AJ/112/2855 Vandalfsen M.L.: J/ApJ/599/1067 Vandame B.: J/A+A/379/436 J/A+A/379/740 J/A+A/384/81 J/A+A/452/119 Vande Putte D.: J/A+A/470/317 Vandehei T.: J/AJ/127/334 J/ApJ/542/257 J/ApJ/552/289 J/ApJ/562/337 J/ApJ/598/597 Vanden Berk D. E.: J/ApJS/122/355 Vanden Berk D.: J/AJ/126/1023

Vanden Berk D.E.: J/AJ/119/2571 J/AJ/121/2308 J/AJ/122/503 J/AJ/123/567 J/AJ/124/1651 J/AJ/125/1711 J/AJ/126/2125 J/AJ/126/2209 J/AJ/126/2579 J/AJ/130/387 J/AJ/130/873 J/AJ/131/84 J/ApJ/633/638 J/ApJ/641/78 J/ApJS/155/257 J/ApJS/163/80 J/ApJS/165/1

Vandenberg D.A.: J/AJ/117/247 J/AJ/127/1227 J/AJ/130/1693 J/AJ/96/1632 J/AJ/97/1360 J/PASP/116/1012

J/PASP/99/739 VANDENBERG D.A.: VI/55 VI/55 Vandenbroucke J.: J/other/APh/26.282 Vandenbussche B.: J/A+A/384/473 J/A+A/390/1033 J/A+A/463/243 J/A+A/475/1159 Vanderriest C.: J/A+AS/114/565 Vanderspek R.: J/other/GCN/1070.1 Vanelle C.: J/A+A/358/77 Vanhollebeke E.: J/A+A/463/243 J/MNRAS/376/313 VanLeeuwen J.: J/AJ/110/2400 Vanmunster T.: J/other/IBVS/5526 Vannier M.: J/A+A/465/469 VANSEVICIUS V: II/196 Vansevicius V.: J/A+A/442/281 J/ApJS/118/177 J/BaltA/7/369 J/PASJ/56/1025 van’t Leven J.: J/A+AS/128/179 VAN’T VEER F.: J/A+A/289/871 van’t Veer F.: J/A+A/311/523 van’t Veer Menneret C.: J/A+A/404/661 Van’t Veer-Menneret C.: J/A+A/405/1095 Vanzella E.: J/A+A/375/1 J/A+A/434/53 J/A+A/449/951 J/A+A/454/423 J/A+A/478/83 J/AJ/122/2190

J/AJ/124/2955 J/ApJ/600/L99 Varadi M.: J/A+A/410/879 J/A+A/430/1049 J/A+A/445/617 J/other/P+SS/53.925 VARELA A.M.: J/A+AS/93/557 Varela A.M.: J/AJ/110/212 Varela J.: J/A+A/393/389 J/A+A/416/475 J/A+A/420/873 J/A+A/470/39 Varendorff M.: J/A+AS/143/145 Varjonen N.: J/A+A/427/769 VARNUM W.B.: I/25 Varricatt W.P.: J/AJ/116/1447 J/AJ/117/2980 J/MNRAS/359/2 J/MNRAS/373/781 Varshalovich D.A.: J/A+A/412/707 VARSHALOVICH D.A.: J/A+AS/104/89 Varshalovich D.A.: J/PAZh/30/430 Vasconcelos M.J.: J/AJ/130/1707 Vashkovjak S.N.: J/A+AS/139/47 Vasilevskis S.: I/11 Vassiliadis E.: J/ApJS/83/87 VASSILIADIS E.: J/ApJS/92/125 Vasundhara R.: J/A+AS/125/399 Vasylyev S.: J/MNRAS/355/1251 Vasyuk V.A.: J/A+A/366/868 J/A+AS/105/503 J/A+AS/140/287 Vaughan A.E.: J/A+A/469/1221 J/AJ/130/2701 V/127 Vaughan A.H.: J/PASP/96/707 Vauglin I.: J/A+A/363/62 J/A+A/387/1 J/A+AS/135/133 VII/234 VII/242 Vavasour J.D.: J/ApJS/90/173 Vaz L.P.R.: J/A+A/304/415 J/A+A/409/1037 J/A+A/461/1065 J/A+A/469/285 VAZ L.P.R.: J/A+AS/101/563 J/A+AS/112/123

Vaz L.P.R.: J/A+AS/115/315 J/A+AS/125/471 J/A+AS/128/251 J/A+AS/130/245 J/A+AS/135/555

J/AJ/127/1117 J/AJ/128/1324 J/AJ/130/2838 Vazdekis A.: J/ApJS/111/203 J/MNRAS/371/703 J/MNRAS/374/664 Vazdekis: J/MNRAS/326/959 VAZQUEZ A.: J/A+A/301/396 Vazquez A.M.: J/MNRAS/359/1524 Vazquez R.: J/A+A/442/917

Vazquez R.A.: J/A+A/371/908 J/A+A/381/884 J/A+A/402/549 J/A+A/409/933 J/A+A/417/961

J/A+A/430/471 J/A+A/432/491 J/A+A/445/493 VAZQUEZ R.A.: J/A+AS/106/339 J/A+AS/111/85

Vazquez R.A.: J/A+AS/116/75 J/A+AS/123/63 J/A+AS/124/13 J/A+AS/137/233 J/MNRAS/367/1441

J/other/NewA/12.461 J/other/RMxAA/39.89 Vedder P.W.: J/AJ/107/751 Vedrenne G.: IX/36 Veen P.M.: J/A+A/384/1012 Vega Beltran J.C.: J/A+A/342/671 J/A+A/360/439 J/A+A/374/394 J/A+AS/131/105 Vega E.I.: J/A+A/471/165 J/A+AS/144/195 VEGA E.I.: J/AJ/105/258 Vega-Beltran J.C.: J/A+A/424/447 Veiga C. H.: J/A+A/425/1107 Veiga C.: J/A+AS/136/445 Veiga C.H.: J/A+A/380/727 J/A+A/400/1095 J/A+A/437/1147 J/A+A/453/349 J/A+A/454/683 VEIGA C.H.: J/A+AS/107/559

Veiga C.H.: J/A+AS/113/557 J/A+AS/120/107 J/A+AS/131/291 J/A+AS/138/247 J/A+AS/139/305

J/A+AS/143/405 Veilleux S.: J/A+A/409/115 J/ApJ/522/113 J/ApJS/143/277 J/ApJS/143/315 VELA VILLAHOZ E.: J/A+AS/103/293 Venemans B.P.: J/A+A/431/793 J/A+A/461/823 Venn K.A.: J/A+A/334/987 J/A+A/367/86 J/A+A/437/467 J/A+A/459/423 J/AJ/127/2723 J/AJ/128/1177

J/AJ/130/2140 Vennes S.: J/MNRAS/335/673 Vennik J.: J/A+AS/117/261 J/A+AS/142/399 J/AN/327/365 Vennik J.A.: J/A+A/296/643 Ventura P.: J/A+A/397/147 J/A+A/418/525 Venturi T.: J/A+A/370/409 J/A+A/379/755 J/A+A/385/39 J/A+A/419/71 J/A+A/423/935 J/A+A/463/519

J/MNRAS/285/898 J/MNRAS/324/1131 Venturini C.C.: J/A+A/459/875 Vera-Villamizar N.: J/AJ/126/1707 Verbunt F.: J/A+A/309/116 J/A+A/382/899 J/A+A/418/509 Verdes-Montenegro L.: J/A+A/411/391 J/A+A/436/443 J/A+A/449/937 J/A+A/462/507 J/A+A/470/505

J/A+A/472/121 Verdugo E.: J/A+AS/137/351 Vergani D.: J/A+A/457/79 J/A+A/463/519 J/A+A/465/1099 Vergani S.: J/A+A/469/503 Vergani S.D.: J/A+A/474/793 Vergely J.L.: J/A+A/411/447 Vergne M.M.: J/A+A/409/933 J/A+A/430/471 J/A+A/462/621 J/A+A/471/165 Verheijen M.A.W.: J/A+A/370/765 J/AJ/112/2471 J/MNRAS/376/125 Verhoelst T.: J/A+A/463/243 Verkhodanov O. V.: J/AZh/72/613 Verkhodanov O.V.: J/ApJ/523/87 J/AZh/71/352 J/AZh/73/835 J/AZh/76/83 J/AZh/77/3 J/AZh/79/589

J/other/BSAO/40.5 J/other/BSAO/49.53 J/other/BSAO/50.115 Verkhodanova N.V.: J/other/BSAO/49.53 Verley S.: J/A+A/436/443 J/A+A/449/937 J/A+A/462/507 J/A+A/470/505 J/A+A/472/121 Verma A.: J/AJ/127/3075 J/AJ/131/2383 J/MNRAS/351/1290 J/MNRAS/379/1343 Verma R.P.: J/A+A/463/175 J/A+A/468/1001 Vermeulen R.: J/AJ/123/1784 Vermeulen R.C.: J/A+A/472/763 J/AJ/130/23 J/AJ/130/2473 J/ApJ/430/467 J/ApJ/609/539 J/ApJS/100/1

J/ApJS/95/345 J/ApJS/98/1 J/ApJS/98/33 VERNER D.A.: J/A+AS/108/287 J/A+AS/109/125 Vernet J.: J/A+A/454/423 J/A+A/465/1099 VERON P.: J/A+A/296/665 Veron P.: J/A+A/414/487 J/A+A/417/515 J/A+A/475/487 J/A+AS/115/97 J/A+AS/135/437 VERON P.: J/MNRAS/260/202 Veron P.: J/MNRAS/275/1102 J/other/Ap/42.1 VERON P.: VII/126A Veron P.: VII/146 VII/166 VII/188 VII/207 VII/215 VII/224 VII/235 VII/248 VII/54 VII/93 Veron-cetty M.-P.: J/A+A/414/487 Veron-Cetty M.-P.: J/A+A/417/515 J/A+A/475/487 J/A+AS/115/97 J/A+AS/135/437 Veron-Cetty M.P.: J/other/Ap/42.1 VERON-CETTY M.P.: VII/126A Veron-Cetty M.P.: VII/146 VII/166 VII/188 VII/207 VII/215 VII/224 VII/235 VII/248 VII/54 VII/93 Veronig A.: J/A+A/447/735 Veronig: J/A+A/390/707 Verrecchia F.: J/A+A/472/705 Verschueren W.: J/A+A/358/553 Verter F.: VII/139 VII/64 Vessey S.J.: J/MNRAS/294/607 Vestergaard M.: J/AJ/132/231 J/ApJ/599/116 J/ApJ/613/682 J/ApJS/134/35 J/ApJS/150/165 Vestrand W. T.: J/ApJS/120/409 Vestrand W.T.: J/AJ/127/2436 J/AJ/128/2965 J/MNRAS/370/435 Vetere L.: J/A+A/473/347 Vetois J.: J/A+A/401/959 Vettolani G.: J/A+A/342/1 J/A+A/379/21 J/A+A/379/35 J/A+A/403/857 J/A+A/428/1043 J/A+A/457/517

J/A+A/457/79 J/A+AS/101/259 J/A+AS/106/1 J/A+AS/123/219 J/A+AS/130/323 J/A+AS/64/247 J/A+AS/95/1 J/MNRAS/296/599

VETTOLANI G.: VII/70A Vettolani G.: VII/71 VII/99 VIII/63 Vettolani P.: J/AJ/128/1974 Viala Y.P.: J/A+A/328/617 VIALE A.: J/A+A/280/365 Viale A.: J/A+AS/31/439 Viallefond F.: J/ApJS/120/247 Viateau B.: I/272 J/A+A/320/652 Vicari A.: J/A+A/357/437 Vicente B.: II/270 J/A+AS/136/307 J/AJ/127/1158 Vicente: J/A+A/476/199 Vicente S.M.: J/A+A/441/195 Vida K.: J/A+A/430/1049 Vidal J.: J/A+AS/125/399 VIDAL-MADJAR A.: J/A+A/299/557 Vidal-Madjar A.: J/A+A/303/137 J/A+A/306/924 J/A+A/385/986 J/A+A/389/149 J/A+A/454/185 VIDAL-MADJAR A.: J/A+AS/109/447 Vidal-Madjar A.: J/A+AS/145/275 Vidal-Sainz J.: II/242 Viehmann T.: J/A+A/433/117 Vieira G.: J/A+AS/134/75 VIEIRA G.: VII/205 Vieira G.G.: J/A+AS/117/335 Vieira K.: J/AJ/127/1158 Vieira Martins R.: J/A+A/380/727 J/A+A/400/1095 J/A+A/425/1107 J/A+A/437/1147 VIEIRA MARTINS R.: J/A+AS/107/559 Vieira Martins R.: J/A+AS/113/557 J/A+AS/117/335 J/A+AS/120/107 J/A+AS/131/291 J/A+AS/136/445

J/A+AS/138/247 J/A+AS/139/305 J/A+AS/143/405 J/AJ/125/2728 Vieira S.L.A.: J/AJ/126/2971 Vienne A.: J/A+A/324/366 J/A+A/380/727 J/A+A/383/296 J/A+A/400/1095 J/A+A/410/343 Vienne A: J/A+A/427/371 Vig S.: J/A+A/463/175 J/A+A/468/1001 Vignali C.: J/A+A/370/900 J/A+A/409/79 J/A+A/432/69 J/A+A/457/501 J/A+A/466/31 J/AJ/124/2351

J/AJ/126/539 J/AJ/130/387 J/ApJ/570/100 J/ApJS/161/21 J/MNRAS/327/771 J/MNRAS/342/575 Vigneau J.: V/13 V/60 Vigotti M.: J/A+A/369/380 J/A+A/379/21 J/A+A/379/35 J/A+A/406/579 J/A+A/435/863 J/A+AS/123/219

J/A+AS/133/129 J/A+AS/139/359 J/AJ/98/419 J/MNRAS/200/747 J/MNRAS/348/857 J/MNRAS/370/1034 Vigouroux G.: J/A+AS/137/269 Vigroux L.: J/A+A/303/137 J/A+A/389/149 J/A+A/454/185 J/A+A/462/81 VIGROUX L.: J/A+AS/109/447 Vigroux L.: J/MNRAS/351/1290 Vikhlinin A.: J/AJ/131/527 J/ApJ/502/558 J/ApJ/594/154 J/ApJ/641/140 J/ApJ/645/955 J/ApJS/161/9 Vikhlinin A.A.: J/ApJS/150/19 Vila B.: J/MNRAS/311/456 Vilardell F.: J/A+A/459/321 Vilas-Boas J.W.S.: J/A+AS/116/21 Vila-Vilaro B.: J/MNRAS/351/1290 Vilchez J.: J/A+AS/128/545 Vilchez J.M.: J/A+A/384/383 J/A+A/397/421 J/A+AS/129/455 J/AJ/110/212 Vilchez-Gomez R.: III/232 J/A+AS/115/219 Vilenius E.: J/A+A/427/769 Villalobos A.: J/AJ/131/2035 Villamariz M.R.: J/A+A/437/467

Villanova S.: J/A+A/407/527 J/A+A/419/149 J/A+A/428/67 J/A+A/434/1107 J/A+A/466/931

J/AJ/130/652 J/MNRAS/367/1441 Villar-Martin M.: J/MNRAS/351/1290 Villasenor J.: J/other/GCN/1070.1

Villata M.: J/A+A/327/61 J/A+A/395/11 J/A+A/402/151 J/A+AS/121/119 J/A+AS/127/445 J/A+AS/129/577 J/A+AS/132/361 J/A+AS/144/481 Vilmer N.: J/A+AS/130/233 VINCENT A.: I/74 VINCENT M.: I/74 Vingerhoets P.: J/A+A/437/705 Vink J.: II/202 J/A+A/304/415 Vink J.S.: J/A+A/439/1107 Vinko J.: J/A+A/465/943 Vinnberg A.: J/AZh/72/22 VINTON G.: J/AJ/107/1453 VIO R.: J/A+AS/112/347 Vio R.: J/A+AS/121/461 Viotti R.: J/A+A/258/423 J/A+AS/136/95 J/A+AS/92/685 VIOTTI R.: J/ApJS/71/983 Viotti R.: VI/48 VI/48 Viotti R.F.: J/A+A/458/225 Virani S.N.: J/AJ/131/2373 Virdefors B.: III/69A Virtanen J.: J/A+A/350/513 J/A+A/374/1049 J/A+A/427/769 VISSER A.E.: J/A+AS/110/419 Visvanathan N.: J/A+AS/118/441 J/ApJS/107/521 Vitale V.: J/A+A/406/L9 J/A+A/410/813 Vitali F.: J/A+A/422/141 J/A+A/469/503 Viti S.: J/A+A/416/655 Vitiello E.: J/MNRAS/264/273 Vitkovskii V.V.: J/AZh/72/613 VITON M.: J/A+AS/107/385 Vitores A.G.: J/A+AS/118/7 J/A+AS/120/323 J/ApJS/105/343 J/ApJS/95/387 Vitrichenko E.A.: J/PAZh/26/525 J/PAZh/27/381 Vittone A. Friedjung M. (eds.): VI/48 Vittone A.: J/A+AS/136/95 Vittone A.A.: J/A+A/367/199 J/A+AS/114/341 Vittorio N.: J/A+AS/133/403 Vitvitska M.: J/ApJ/598/260 Vivas A.K.: II/266 J/AJ/127/1158 J/AJ/129/907 J/AJ/132/714 Vivekananda Rao P.: J/A+A/297/359 Vlasuyk V.V.: J/A+A/376/775 J/A+A/422/925

Vlasyuk V.V.: J/A+A/371/79 J/AZh/81/925 J/other/AISAO/32.31 J/PAZh/25/18 J/PAZh/27/83

J/PAZh/27/942 Voelk H.J.: J/A+A/390/39 J/A+A/406/L9 J/A+A/410/813 VOGEL S.: J/A+A/294/377 Vogeley M.S.: J/AJ/106/1273 J/AJ/127/1811 J/ApJS/121/287 Voges W.: II/203A IX/10A IX/11 IX/29 IX/32 IX/35 J/A+A/323/739 J/A+A/326/211 J/A+A/330/108

J/A+A/335/467 J/A+A/347/47 J/A+A/356/445 J/A+A/406/535 J/A+A/414/487 J/A+A/420/61 J/A+A/423/449 J/A+A/425/367 J/A+A/449/425 J/A+AS/109/147

VOGES W.: J/A+AS/110/469

Voges W.: J/A+AS/127/145 J/A+AS/127/251 J/A+AS/129/87 J/A+AS/132/155 J/A+AS/132/341 J/A+AS/134/221 J/A+AS/135/319 J/AJ/122/503 J/AJ/123/567 J/AJ/123/578 J/AJ/126/1720 J/AJ/126/2209 J/AJ/126/2579 J/AJ/126/2740 J/AJ/127/1811 J/AJ/129/2542 J/AJ/130/968 J/ApJ/514/148 J/ApJ/595/1206 J/ApJ/616/1284 J/ApJ/617/192 J/ApJS/117/319 J/ApJS/129/435 J/ApJS/140/239 J/ApJS/146/267 J/ApJS/148/243 J/ApJS/149/29 J/ApJS/162/304 J/MNRAS/274/1194 J/MNRAS/274/75 J/MNRAS/281/799 J/MNRAS/301/881

Vogler A.: J/A+A/305/74 Vogler F.: J/A+A/447/735 VOGT N.: II/200 Vogt N.: J/A+A/463/1053 J/A+AS/130/455 J/A+AS/93/463 VOGT N.: VII/5A Vogt N.P.: J/AJ/113/1197 J/AJ/113/22 J/AJ/117/1668 J/AJ/127/3273 J/ApJ/481/673 J/ApJ/571/136

J/ApJ/599/1006 J/ApJS/142/1 J/ApJS/157/175 J/ApJS/159/41 Vogt S.S.: J/ApJ/634/625 J/ApJS/141/503 J/ApJS/152/261 Voigt B.: J/other/APh/26.282 Voit G.M.: J/ApJ/569/689 Volchkov A.A.: III/182 Volk K.: J/A+A/368/1006 J/ApJS/152/201 Vollmer B.: J/A+A/418/1 Volonteri M.: J/A+AS/145/111 Voloshina I.B.: III/208 J/AZh/78/408 J/AZh/79/1087 Volp A.W.: J/A+A/352/228 Volten H.: J/A+AS/116/257 von Braun K.: J/AJ/123/279 VON DER HEIDE J.: I/62C von Hippel T.: J/AJ/128/1825 J/AJ/131/582 J/ApJ/562/528 J/ApJ/632/266 von Hippel T.A.: J/AJ/104/1765 von Montigny C.: J/ApJS/101/259 Von Montigny C.: J/ApJS/107/227 J/ApJS/123/79 von Montigny C.: J/ApJS/94/551 von Zeipel M.H.: I/278 Vondrak J.: J/A+A/446/583 J/A+A/463/783 Voronin A.: J/A+A/430/1139 Voronkov M.A.: J/A+AS/134/115 J/MNRAS/367/1609 VORONTSOV-VELYAMINOV B.A.: VII/100 Vorontsov-Velyaminov B.A.: VII/236 VII/62A Voroshilov V.I.: III/230 VOS N.: II/200 Voss R.: J/A+A/447/71 J/A+A/468/49 Votruba V.: J/A+A/419/607 Vozyakova O.V.: J/A+AS/143/211 V.P.: J/AZh/77/420 Vracko Y.A.: J/AJ/106/2058 Vrba F.J.: J/AJ/103/638 J/AJ/105/1571 VRBA F.J.: J/AJ/105/258 Vrba F.J.: J/AJ/114/2822 J/AJ/117/2941 J/AJ/127/3553 J/ApJ/446/115 J/BaltA/12/323 J/BaltA/13/1

J/BaltA/15/483Short J/BaltA/7/369 Vreux J.-M.: J/A+A/418/841 J/A+A/454/1047 J/A+A/477/593 J/MNRAS/372/661 Vrielmann S.: II/202 Vrtilek S.D.: J/ApJ/412/479 J/ApJS/160/450 vsini and X-ray luminosities relationships. Cutispoto G.: J/A+A/384/491 Vu D.T.: J/A+AS/125/399 Vujnovic V.: J/A+A/388/704 V.V.: I/275 VYSSOTSKY A.N. et al.: III/13

Wachter S.: J/ApJ/565/511 J/ApJ/586/1356 J/ApJ/615/887 J/PASP/109/977 Wackerling L.R.: III/17B Wackermann R.: J/A+A/396/513 Wada K.: J/PASJ/52/631 Wadadekar Y.: J/A+A/416/35 J/AJ/118/1435 Waddell P.: J/AJ/121/2308 J/AJ/121/2331 J/AJ/123/567 Waddington I.: J/MNRAS/317/801 J/MNRAS/379/1343 Wade C.M.: J/ApJS/13/65 Wade G.: J/A+A/314/491 J/A+A/347/164

Wade  G.A.:  J/A+A/335/973 J/A+A/361/991 J/A+A/374/265 J/A+A/458/569 J/A+A/475/1053

J/MNRAS/376/361 Wade R.A.: J/AJ/126/1455 Waelde E.: II/202 J/A+A/304/415 WAELKENS C.: J/A+A/286/136

Waelkens C.: J/A+A/307/869 J/A+A/310/164 J/A+A/311/873 J/A+A/336/263 J/A+A/390/1033

J/A+A/415/241 J/A+A/463/243 J/A+A/465/965 Wagg J.: J/MNRAS/372/1621 Wagner R. M.: J/ApJ/510/659 Wagner R.M.: J/A+A/312/439 J/A+A/369/57 J/A+A/410/813 J/A+A/434/1107 J/ApJS/115/185 Wagner S.: J/A+A/323/707 J/A+A/422/907 J/A+A/466/41 Wagner S.J.: J/A+A/398/49 J/ApJS/115/185 Wagner W.: J/other/APh/26.282 Wahlgren G.M.: J/A+A/418/1073 J/A+A/465/357 Wahnon P.: J/A+AS/109/201 Wainscoat R.J.: J/AJ/112/2471 J/ApJ/602/571 Wakamatsu K.: J/A+A/413/843 J/AJ/112/2408 J/MNRAS/354/980 J/MNRAS/368/534 Wakamatsu K.-I.: J/MNRAS/316/326 Wake D.: J/MNRAS/372/425 J/MNRAS/375/68 Wakker B.P.: J/ApJS/136/463 J/ApJS/146/1 J/ApJS/146/125 Walborn N.R.: III/115 III/188 III/195 J/A+A/436/117 J/AJ/103/1205 J/AJ/118/1684 J/MNRAS/375/145

V/116 Walck C.: J/other/APh/26.282 Waldhausen S.: J/A+A/355/138 Waldmann H.: J/other/APh/26.282 Waldram E.: J/MNRAS/341/1066 J/MNRAS/360/340 WALDRAM E.M.: J/A+AS/110/419 Waldram E.M.: J/MNRAS/246/110 J/MNRAS/276/614 J/MNRAS/282/779 J/MNRAS/298/637 J/MNRAS/307/293

J/MNRAS/342/915 J/MNRAS/354/485 J/MNRAS/379/1442 VIII/31 Waldthausen H. et al: VIII/43 Walker A.: J/AJ/119/1824 J/AJ/123/3216 J/AJ/128/1167 J/MNRAS/246/54 Walker A.R.: II/179 III/185 J/A+A/371/579 J/AJ/105/527 J/AJ/108/555 J/AJ/110/638 J/AJ/112/2408

J/AJ/127/958 J/AJ/98/2086 J/ApJ/365/471 J/ApJ/562/303 J/ApJS/71/25 J/MNRAS/224/61 J/MNRAS/300/665 J/PASP/104/1063 J/PASP/106/828

Walker C.E.: J/AJ/112/62 Walker D.W. (Joint IRAS Science W.G.): VII/73 Walker G.: J/ApJ/549/578 J/other/IBVS/5526 Walker G.A.H.: III/185 J/A+A/312/879 J/ApJ/459/278 Walker H.J.: J/ApJS/151/299 WALKER I.: J/AJ/106/591 Walker M.A.: J/MNRAS/354/305 J/MNRAS/371/898 Walker M.G.: J/AJ/131/2114 J/AJ/131/912 Walker R.G.: J/AJ/110/275 Walker R.L.: J/AJ/105/1571 J/AJ/106/2058 Walker R.N.F.: J/MNRAS/362/689 Walkowicz L.M.: J/AJ/123/3409 J/AJ/128/426 Wall J.: J/MNRAS/346/627 Wall J.V.: J/A+A/386/97 J/A+A/399/469 J/MNRAS/200/747 J/MNRAS/216/173 Wallace B.J.: J/ApJS/107/239 Wallace L.: III/240 J/ApJS/103/235 J/ApJS/107/312 J/PASP/107/1042 Wallace P.T.: J/A+A/400/1145 J/A+A/406/1135 J/A+A/450/855 J/A+A/459/981 WALLENQUIST A.: II/101 Waller W.H.: J/AJ/111/1128 J/AJ/113/1011 J/ApJS/63/821 Wallerstein G.: J/AJ/118/1245 J/AJ/127/373 J/AJ/127/980 J/PASP/111/335 Wallyn P.: J/ApJS/127/79 Walmsley C.M.: J/A+A/321/293 J/A+A/432/921 Walraven J.H.: J/A+AS/137/113 Walraven T.: J/A+AS/137/113 Walsh A.: J/MNRAS/367/1609

Walsh A.J.: J/A+A/410/597 J/A+A/453/1003 J/A+A/466/989 J/MNRAS/291/261 J/MNRAS/301/640

J/MNRAS/363/405 J/MNRAS/368/1223 Walsh B.Y.: J/A+A/411/447 Walsh J.R.: J/A+AS/125/289 Walter F.: J/A+A/459/875 J/AJ/120/1794 J/ApJ/633/871 J/other/IBVS/5731 Walter F.M.: J/A+A/468/1039 J/A+AS/146/323 J/AJ/116/941 J/AJ/128/2316 J/ApJS/165/256 WALTER H.: I/169 Walter H.G.: J/A+AS/122/529 V/67 Walter M.: J/other/APh/26.282 WALTER R.: J/A+A/274/105 Walter R.: J/A+A/307/33 J/ApJ/607/L33 J/ApJ/636/765 J/ApJS/170/175 Walterbos R.A.M.: J/A+A/406/829 J/A+AS/61/451 J/A+AS/92/625 J/ApJ/613/914 Walth G.L.: J/AJ/129/393 Waltman E.B.: J/ApJS/136/265 WALTON MAYALL M.: III/138 Walton N.A.: J/A+A/424/545 J/A+A/472/805 Wambsganss J.: J/A+A/440/53 Wamsteker W.: II/65 J/A+A/369/57 J/A+A/373/730 J/A+AS/121/461 J/ApJS/99/223 Wandel A.: J/ApJ/613/682 Wanders I.: J/ApJ/510/659 WANG G.: VII/169 Wang H.: J/ApJS/141/157 Wang H.Q.: J/A+A/430/327 Wang H.X.: J/MNRAS/362/1189 Wang J.: I/258 J/A+A/380/665 J/A+AS/128/315 J/AJ/122/2587 J/ApJS/139/369 WANG J.C.: J/A+AS/106/361 Wang Jiancheng: J/A+AS/129/569 Wang J.J.: I/257B I/257B J/A+AS/142/373 J/A+AS/145/223 J/AJ/125/1397 Wang J.-M.: J/ApJ/618/L83 Wang J.S.: J/A+AS/127/181 Wang J.-S.: J/other/PASA/17.92 Wang J.S.: J/other/PPMtO/14.185 Wang J.X.: J/A+A/451/457 J/AJ/127/213 Wang J.-X.: J/ApJ/645/856 Wang J.X.: J/ApJS/155/271 J/ApJS/155/73 Wang L.: J/A+A/437/789 J/ApJ/645/488 Wang M.: J/ApJS/141/157 Wang N.: J/MNRAS/328/855 J/MNRAS/362/1189 Wang Q.D.: J/AJ/129/31 J/ApJ/554/202 J/ApJ/598/969 J/ApJ/611/821 J/ApJS/165/173 J/MNRAS/362/1065

J/MNRAS/371/147 J/MNRAS/371/38 J/MNRAS/376/960 Wang R.: J/A+A/383/1062 J/A+AS/118/239 J/ApJ/645/890 J/other/PBeiO/17.120 Wang S.: J/A+AS/99/545 VIII/44 Wang S.H.: J/A+A/454/379 J/MNRAS/376/1707 Wang T.: J/A+A/309/81 Wang T.-G.: J/A+AS/121/461 J/ApJ/645/856 J/other/ChJAA/2.501 Wang X.: J/AJ/131/2114 J/AJ/131/2245 J/ApJ/645/488 Wang X.-H.: J/A+A/387/129 Wang Y.: J/A+A/461/197 J/AJ/116/2081 Wang Y.-R.: J/other/APh/26.282 Wang Z.: J/A+AS/118/1 J/A+AS/136/1 J/AJ/131/527 Wang Z.Z.: J/A+A/430/327 J/other/PBeiO/32.87 Wannier P.G.: J/AJ/124/2164 Ward M.J.: J/A+A/476/1191 J/MNRAS/357/401 Ward-Thompson D.: J/A+A/372/L41 Wargau W.: J/MNRAS/274/1194 J/MNRAS/289/729 WARGAU W.F.: J/A+AS/107/277 Wargau W.F.: J/A+AS/125/71 Wark R.: J/other/PASAu/10.310 WARK R.M.: J/MNRAS/251/330 Warner B.: J/A+A/465/331 Warner P.J.: J/MNRAS/246/110 J/MNRAS/276/614 J/MNRAS/282/779 WARNER P.J.: VIII/21 VIII/22 VIII/23 VIII/24 VIII/25 Warner P.J.: VIII/31 Warnock A.: II/115A Warnock A. III: II/81 WARNOCK A.: VII/63 Warren B.E.: J/AJ/128/16 Warren: J/MNRAS/325/1002 Warren Jr. W.H.: IV/12 V/102 V/105 V/109 Warren Jr W.H.: V/50

Warren S.J.: J/ApJ/462/614 J/ApJS/104/185 J/MNRAS/328/150 J/MNRAS/352/285 VII/148 VII/182

VII/182 Warren W.H.: I/71 II/36 III/89 Warren W.H. Jr.: I/171 I/172 I/201 V/95 Warren W.H.: VI/29A VI/34 Warwick R.S.: J/A+A/476/1191 J/MNRAS/351/31 J/MNRAS/357/401 J/MNRAS/370/25 Washuettl A: J/A+A/430/1049 Washuettl A.: J/A+AS/137/369 J/A+AS/142/275 Waskett T.J.: J/MNRAS/350/785 WASSERMAN L.H.: J/AJ/107/1814 Wasserman L.H.: J/AJ/109/1363 Watanabe C.: J/ApJ/610/128 Watanabe E.: II/192 J/PASJ/57/13 Watanabe M.: VII/147 Watanabe S.: J/ApJS/112/245 Watanabe T.: J/PASJ/52/631 WATANABE T.;: VII/140 Watase Y.: J/MNRAS/330/137 Waterhouse E.: J/AJ/119/2214 Waters L.: J/A+A/390/1033 WATERS L.B.F.M.: J/A+A/290/609

Waters L.B.F.M.: J/A+A/307/869 J/A+A/310/893 J/A+A/323/469 J/A+A/352/228 J/A+A/382/184

J/A+A/384/473 J/A+AS/125/419 Watson A.: J/AJ/113/1001 Watson A. M.: J/PASP/108/682 Watson A.M.: J/AJ/117/1700 J/AJ/129/393 Watson C.: J/A+A/447/921 J/ApJ/635/560 J/ApJS/154/553 Watson D.: J/A+A/385/377 J/PASP/118/1690 Watson F.G.: J/AJ/123/2490 J/MNRAS/300/417 J/MNRAS/369/1131 Watson M.: J/A+A/474/473 J/MNRAS/372/741 Watson M.G.: J/A+A/428/383 J/A+A/469/27 J/A+A/476/1191 J/MNRAS/274/1194 J/MNRAS/311/456

J/MNRAS/351/31 VII/19 Watson R.A.: J/MNRAS/341/1066 J/MNRAS/360/340 Watts C.B.: I/100A I/100A VI/122 Waugh M.: J/AJ/124/1954 J/AJ/128/16 J/MNRAS/337/641 Way M.J.: J/AJ/112/36 J/AJ/130/2012 Wayman P.A.: J/MNRAS/259/563 Weaver G.: J/AJ/114/2822 Weaver H.: VIII/11 Weaver K.A.: J/ApJ/602/231 J/ApJS/96/303 Weaver W.B.: III/181 J/AJ/114/2822 J/ApJ/446/300 WEAVER W.B.: J/ApJS/78/239 Webb D.: VII/192 Webb J.: J/A+AS/134/401 J/ApJS/107/175 Webb J.K.: J/ApJ/498/77 J/ApJ/559/654 J/MNRAS/345/609 J/other/PASA/19.455 Webb N.A.: J/A+A/400/521 J/A+A/445/155 J/A+A/476/1191 Webb T.M.A.: J/AJ/117/1313 Webb W.: J/ApJS/125/349 Webbink R.F.: J/AJ/125/2239 J/AJ/127/477 J/other/IBVS/5298 V/110 V/123A V/94 VII/151 Weber M.: J/A+A/430/1049 J/A+AS/137/369 J/A+AS/142/275

Webster R.L.: J/AJ/106/848 J/AJ/121/2381 J/AJ/123/873 J/AJ/124/1954 J/AJ/124/690 J/AJ/128/16

J/ApJS/165/307 J/MNRAS/337/641 VIII/34 Webster: VIII/73 Webster Z.T.: J/AJ/116/1757 Wechsler R.H.: J/AJ/130/968 Weedman D.: J/ApJ/634/128 J/ApJ/634/L1 Wegner G.: J/A+A/407/423 J/A+AS/140/327 J/A+AS/141/449 J/AJ/105/1251 J/AJ/105/788 J/AJ/109/2368

J/AJ/113/1197 J/AJ/113/22 J/AJ/116/2108 J/AJ/117/1668 J/AJ/117/2039 J/AJ/122/2893 J/AJ/124/1266 J/AJ/125/2307 J/AJ/126/2268 J/AJ/130/496 J/ApJ/470/724 J/ApJS/106/1 J/ApJS/140/265 J/MNRAS/292/499 J/MNRAS/305/259 J/MNRAS/321/277 VII/176A

Wegner G.A.: J/AJ/120/2338 J/AJ/128/1558 Wegner: J/A+A/395/753 Wegner M.: J/A+A/466/255 Wegner R.: J/ApJ/439/983 Wegner W.: II/175 J/AcA/43/209 J/AN/324/219 J/BaltA/11/1 Wehlau A.: J/AJ/117/286 Wehlau W.H.: J/ApJ/459/278 Wehmeyer R.: III/205 J/AN/324/437 Wei J.-Y.: J/A+A/456/261 Wei J.Y.: J/A+AS/139/575 Wei J.-Y.: J/ApJ/510/659 Wei J.Y.: J/ApJ/590/73 J/ApJS/115/185 Wei L.H.: V/116 Wei W.-R.: J/PASJ/52/997 Wei Y.: J/A+A/426/503 Weidenschilling S.: J/PASP/118/1690 Weidenspointner G.: J/ApJS/156/69 Weidner C.: J/A+A/402/571 Weigelt G.: J/A+A/287/769 J/A+A/288/413 J/A+A/385/87 J/A+A/394/617 J/A+A/409/245 J/A+A/422/627

J/A+A/448/1075 J/A+A/466/1099 Weikard H.: J/A+A/309/581 Weiland J.L.: J/A+AS/127/1 J/ApJS/148/97 Weilbacher P.: J/MNRAS/372/425 Weilbacher P.M.: J/A+A/358/819 Weinberg D.H.: J/AJ/126/2209 J/ApJ/633/844 J/ApJS/163/80 Weinberg J.L.: J/A+AS/127/1 Weinberg M.: II/246 J/AJ/128/245 WEINBERG M.D.: J/ApJ/411/188

Weinberger R.: J/A+A/286/17 J/A+AS/110/269 J/A+AS/117/1 J/A+AS/117/369 J/A+AS/126/247 J/A+AS/137/293 WEINBERGER R.: J/A+AS/93/553 Weinberger R.: J/AJ/107/2101 Weiner B.J.: J/ApJ/599/1006 J/ApJ/617/718 J/ApJ/625/6 J/ApJ/643/680 J/ApJS/142/1 J/ApJS/157/175

J/ApJS/159/41 Weinstein M.A.: J/ApJS/155/257 Weis E.W.: J/AJ/105/1962 Weisberg J. M.: J/ApJS/121/171 Weiss A.: J/A+A/361/1023 J/A+A/396/503 WEISS G.: III/93 Weiss W.W.: J/A+A/373/625 J/A+AS/105/125 J/A+AS/123/93 J/MNRAS/357/345 Weisskopf M.C.: J/ApJ/615/897 Weistrop D.: II/87 J/AJ/125/1696 J/AJ/127/1360 Weitenbeck A.J.: J/AcA/49/59 J/AcA/54/87 J/AJ/114/222 Weitz K. A.: J/ApJS/121/171 Welch D.: J/AJ/112/2408 WELCH D.L.: J/AJ/105/938 J/AJ/106/1508 Welch D.L.: J/AJ/107/2093 J/AJ/114/326 J/AJ/117/1313 J/AJ/127/334 J/ApJ/454/788 J/ApJ/542/257

J/ApJ/552/289 J/ApJ/562/337 J/ApJS/90/83 Welch D.L. (the Macho Collaboration): J/ApJ/598/597 Welch G.A.: J/AJ/113/2104 Wells A.A.: J/A+A/448/L9 Wells L.A.: J/AJ/112/2408 Welsh B.Y.: J/A+A/346/785 J/AJ/130/825 J/ApJ/619/L95 J/ApJS/164/38 Weltevrede P.: J/A+A/445/243 J/A+A/469/607 Welty A.D.: J/ApJS/123/283 Welty D.E.: II/149 J/ApJS/147/61 J/ApJS/151/313 Wen W.: J/other/ChJAA/6.435 WENDEROTH E.: J/AJ/107/230 Wenderoth E.: J/AJ/99/1831 WENDKER H.J.: II/112 II/129 II/147 Wendker H.J.: II/199A J/A+A/241/551 WENDKER H.J.: J/A+A/281/161 Wendker H.J.: J/A+A/299/39 J/A+AS/126/509 J/AN/321/81 Wendlandt H.U.: VIII/47 Wendt C.: J/other/APh/26.282 Wendt G.: J/A+A/467/585 Wenk R.: J/AJ/116/102 Wenz M.T.: J/PASP/110/1012 WENZEL W.: VI/17 Wenzel W.: VI/35 VI/38 VI/67 Werk J.: J/ApJS/165/307 Werk J.K.: J/AJ/130/2571 J/AJ/130/2584 J/AJ/130/496 Werner K.: J/A+A/470/317 Werner M.: J/AJ/127/2659 J/ApJS/149/365 J/ApJS/154/155 Werner M.W.: J/AJ/125/2411 J/ApJ/636/1098 Wertheimer J.G.: J/ApJS/169/430 Weselak T.: J/A+A/414/949 J/MNRAS/358/563 WESEMAEL F.: J/AJ/105/35 Wesselink A.J.: J/MNRAS/121/337 Wesselius. P.: J/A+A/379/564 WESSELIUS P.R.: II/114 II/148 Wesselius P.R.: II/97 J/A+A/365/110 Wesselius P.R..: J/A+A/378/116 WESSELIUS P.R.: J/A+AS/100/473 Wesson K.: J/AJ/121/225 West A.: J/ApJ/633/844 West A.A.: J/AJ/128/426 J/AJ/130/1871 J/AJ/131/1674 West D.: J/other/IBVS/5526 West M.J.: J/A+A/451/1159 J/A+A/459/679 J/AJ/127/24 J/ApJ/613/279 J/ApJS/153/223 J/ApJS/154/509 West R.G.: J/A+A/467/785 J/MNRAS/274/1165 J/MNRAS/372/1117 J/MNRAS/379/647 J/MNRAS/379/816 West R.M.: VI/30A Westera P.: J/A+A/423/133 Westergaard N.J.: J/A+A/427/87 Westerhout G.: VIII/47 VIII/64 Westerlund B.: II/221A WESTERLUND B.: J/A+A/293/75 Westerlund B.E.: J/A+A/276/25 WESTERLUND B.E.: J/A+AS/110/533 Westerlund B.E.: J/A+AS/4/51 WESTERLUND B.E.: J/A+AS/76/101 Westerlund B.E.: J/A+AS/97/603 J/MNRAS/130/45 WESTERLUND B.E. STEENMAN H.C.: J/A+AS/105/311 Westin J.: J/ApJ/530/783 Westpfahl D.J.: J/PASP/111/1169 Westphal J. A.: J/PASP/108/682 Westphal J.: J/AJ/113/1001 Westphal J.A.: J/AJ/117/1700 Wetterer C.J.: J/AJ/112/742 Weymann R.J.: J/AJ/116/941 J/AJ/117/17 J/ApJ/546/694 J/ApJ/636/631 W.G.L.: J/A+AS/142/25 Whalen D.J.: J/AJ/131/1948 Wheatley J.M.: J/AJ/130/825 Wheatley P.J.: J/A+A/445/155 J/A+A/467/785 J/A+A/476/1191 J/MNRAS/372/1117 J/MNRAS/379/647

J/MNRAS/379/816 Wheaton W. A.: J/ApJS/127/79 Wheaton W.A.: II/246 Wheelock S.: II/246 Whitaker K.E.: J/MNRAS/362/1065 White B.: J/ApJ/620/1010 White B.A.: J/ApJ/633/871 White D.A.: J/MNRAS/292/419 White G.: J/A+AS/125/399 White G.J.: J/MNRAS/375/115

White G.L.: J/A+AS/111/311 J/A+AS/120/77 J/A+AS/121/321 J/A+AS/127/119 J/A+AS/130/421

J/MNRAS/227/607 J/MNRAS/355/44 J/other/PASAu/10.140 J/other/PASAu/10.310 J/other/SerAJ/166.13Short WHITE G.L.: VIII/45 White J.: II/156A White N.E.: IX/12 IX/31 J/A+AS/109/9 VI/43 White R.: J/ApJ/621/53 White R.A.: J/AJ/118/2014 J/ApJS/80/501 WHITE R.A.: VIII/29 White R.E.: II/37 III/89 VII/101A WHITE R.E.: VII/31B VII/44B White R.L.: J/A+A/356/445 J/AJ/112/407 J/AJ/118/1684 J/AJ/121/1872 J/AJ/121/31 J/AJ/127/2565

J/AJ/128/1990 J/AJ/129/348 J/AJ/130/156 J/AJ/130/586 J/AJ/131/666 J/ApJ/461/127 J/ApJ/607/60 J/ApJ/612/202 J/ApJ/616/1284 J/ApJS/126/133 J/ApJS/135/227 J/ApJS/143/1 J/ApJS/150/1 J/ApJS/74/181 J/ApJS/91/347 VIII/13 VIII/17 VIII/48 VIII/51 VIII/59 VIII/71

White S.: J/MNRAS/371/1777 White S.D.M.: J/A+A/427/397 J/A+A/444/365 VII/221 WHITELOCK P.: J/MNRAS/267/711 Whitelock P.A.: J/A+A/406/51 J/A+A/452/503 J/A+A/463/703 J/A+A/465/481 WHITELOCK P.A.: J/A+AS/93/383 Whitelock P.A.: J/MNRAS/317/460 J/MNRAS/369/751 J/MNRAS/369/783 J/MNRAS/376/313 Whiteoak J.B.: J/A+A/305/960 VII/216 Whiteoak J.B.Z.: J/A+AS/118/329 Whitford A.E.: J/ApJ/353/494 Whiting A.B.: J/ApJ/622/217 Whitlock S.: J/MNRAS/345/949 Whitmore B.: J/PASP/117/37 Whitmore B.C.: J/AJ/109/960 J/AJ/114/2381 J/AJ/124/1418 J/ApJ/513/733 J/ApJ/561/727 VII/72 Whitney B.A.: J/ApJ/635/560 Whitney J.H.: J/AJ/108/1350 Whittet D.C.B.: J/ApJ/623/897 J/other/SAAOC/15.40 J/other/SAAOC/15.65 J/other/SAAOC/15.73 Whittle M.: J/MNRAS/357/1337 Whyte L.: J/A+A/421/595 W.I.: V/111 Wicenec A.: I/207 Wicenec A: I/259 Wicenec A.: J/A+A/373/576 J/A+A/408/1029 J/A+A/452/119 J/A+AS/137/75 J/A+AS/137/83

Wichmann R.: J/A+A/297/391 J/A+A/312/439 J/A+A/319/184 J/A+A/320/185 J/A+A/326/211

J/A+A/349/619 J/A+AS/114/109 J/A+AS/123/329 J/A+AS/137/75 Widicus Weaver S.L.: J/ApJS/158/188 Wiedner C.A.: J/A+A/390/39 J/A+A/406/L9 J/A+A/410/813

Wielebinski R.: J/A+A/418/1 J/A+A/448/411 J/A+AS/106/303 J/A+AS/109/147 J/A+AS/111/311 J/A+AS/114/21 J/A+AS/120/77 J/A+AS/121/321 J/A+AS/127/119 J/A+AS/135/571 J/A+AS/147/195 J/A+AS/75/67

WIELEBINSKI R.: J/ApJS/93/145 WIELEN R.: I/159A Wielen R.: I/264 I/286 J/A+AS/146/251 Wieprecht E.: J/A+A/390/1033 Wiercigroch A.: J/AJ/107/751 J/ApJS/94/127 Wieringa M.H.: J/A+A/457/517 WIERINGA M.H.: J/A+AS/93/399 Wieringa M.H.: J/AJ/130/1358 J/MNRAS/325/1411 VIII/19 VIII/63 Wiesbusch C.H.: J/other/APh/26.282 WIESE K.: III/155 Wiese W.L.: VI/72 VI/72 Wiesenfeld L.: J/A+A/472/1029 Wietfeldt R.: J/PASJ/52/997 Wiggs M.S.: J/AJ/120/2747 Wijers R.A.M.J.: J/ApJ/634/L89 Wik D.: J/ApJ/618/123 Wikstroem G.: J/other/APh/26.282 Wiktorowicz S.: J/ApJ/636/1020 Wil van Breugel: VIII/69A Wilcox R.E.: J/ApJ/586/1356 Wild P.: VII/4A Wild V.: J/MNRAS/371/495 Wild W.: J/A+AS/122/271 Wildeman K.J.: II/97 Wilden B.: J/ApJS/140/143 Wilds R.P.: J/A+AS/125/399 Wilgenbus D.: J/A+A/356/1010 Wilhelm K.: J/A+A/375/591 J/A+AS/126/281 Wilhelm R.: J/A+A/431/143 J/AJ/108/268 J/AJ/112/1188 J/AJ/117/2308 J/AJ/117/2329 J/AJ/123/3216

J/AJ/126/1362 J/AJ/130/1097 J/ApJS/103/433 J/MNRAS/320/451 Wilhite B.C.: J/AJ/131/1923 J/ApJ/633/638 J/ApJ/641/78 Wilke K.: J/A+A/401/873 Wilkes B. J.: J/ApJ/510/659 Wilkes B.: J/A+A/371/79 J/AJ/129/2074 J/ApJS/92/53 Wilkes B.J.: J/A+A/369/57 J/ApJ/618/123 J/ApJ/644/829 J/ApJ/645/955 J/ApJS/150/19 J/ApJS/95/1 Wilking B.A.: J/AJ/127/1131 J/AJ/130/1733 WILKING B.A.: J/ApJ/397/520 Wilkinson A.: J/MNRAS/341/1066 Wilkinson E.: J/ApJS/146/1 Wilkinson M.I.: J/AJ/131/895 J/MNRAS/330/792 J/MNRAS/369/120 J/MNRAS/369/97 Wilkinson P.: J/A+A/472/763 Wilkinson P.N.: J/A+A/472/805 J/A+A/474/1093 J/ApJS/100/1 J/ApJS/95/345 J/ApJS/98/1 J/ApJS/98/33

J/ApJS/99/297 J/MNRAS/358/1159 J/MNRAS/368/1411 J/MNRAS/376/371 VIII/60 VIII/72 Wilkinson S.: J/MNRAS/330/137 WILL J.M.: J/A+A/301/396 Will J.-M.: J/A+A/328/167 J/A+A/369/511 J/A+A/381/884 WILL J.-M.: J/A+AS/111/85 J/A+AS/112/367 Will J.-M.: J/A+AS/123/455 Willacy K.: J/A+A/330/676 J/A+AS/121/139 William I. Hartkopf: I/281 I/301 Williamon R.M.: J/AJ/128/1319 J/AJ/129/2798 Williams B.F.: J/AJ/124/1651 J/AJ/129/2663 J/ApJ/559/851 J/ApJ/609/735 J/ApJ/610/247 Williams C.C.: J/AJ/127/3553 Williams D.R.: J/other/APh/26.282 Williams D.R.W.: VIII/11 Williams E.C.: J/AJ/127/1602 Williams G.A. (1981): VII/171 Williams G.G.: J/A+A/399/365 Williams G.V.: II/238 IV/29 J/A+AS/115/295 Williams I.P.: J/A+AS/121/65 J/A+AS/130/77 WILLIAMS J.: J/AJ/107/679 Williams J.G.: J/AJ/110/1420 Williams J.P.: J/ApJ/428/693 Williams K.A.: J/AJ/131/582 Williams K.L.: J/AJ/109/543 Williams O.R.: J/A+AS/134/287 J/A+AS/143/145 Williams P.M.: J/A+A/357/572 J/ApJS/149/123 Williams P.R.: J/AJ/128/2965 Williams R.: J/ApJS/149/157 J/MNRAS/362/1065 Williams R.E.: J/A+A/459/875 J/AJ/112/1335 J/AJ/119/486 J/AJ/120/2747 J/AJ/130/1358 J/AJ/99/1079 Williams R.J.: J/AJ/124/3042 Williams S.J.: J/A+A/417/115 Williams T.B.: J/AJ/107/2067 J/AJ/110/1699 J/AJ/112/2408 J/AJ/113/1026 J/AJ/119/1268 J/AJ/131/2089 Williamson R.: J/A+A/317/689 Willick J.A.: J/ApJS/109/333 VII/198 Williger G.M.: J/AJ/112/2408 J/ApJ/614/75 J/ApJ/636/631 Willis D.R.: J/A+A/467/585 J/ApJ/636/765 Willis: J/MNRAS/325/1002 Willis J.P.: J/AJ/134/1103 J/MNRAS/382/279 Willis T.: J/ApJS/94/551 Willis T.D.: J/ApJS/101/259 J/ApJS/107/227 Willman B.: J/AJ/121/1872 Willmarth D.: J/ApJS/162/207 J/PASP/111/1233 Willmarth D.W.: J/ApJ/521/682 Willmer C.N.A.: J/AJ/105/1637 J/AJ/116/1 J/AJ/118/1131 J/AJ/123/1200 J/AJ/125/2307 J/AJ/126/1750

J/AJ/126/2268 J/AJ/127/3121 J/AJ/128/2642 J/ApJ/571/136 J/ApJ/599/1006 J/ApJ/625/6 J/ApJS/142/1 J/ApJS/157/175 J/ApJS/159/41 VII/200

Willner S.P.: J/A+A/403/493 J/AJ/131/527 J/ApJ/640/167 J/ApJ/642/126 J/ApJS/154/155 J/ApJS/95/1 Willott C.: J/AJ/129/2074 Willott C.J.: J/ApJ/633/630 J/MNRAS/343/293 J/MNRAS/351/1290 J/MNRAS/355/97 J/MNRAS/358/333

J/MNRAS/372/1621 J/MNRAS/373/1531 Wills B.J.: J/AJ/127/156 WILLS D.: VII/53 VIII/4 Willson L.A.: J/AJ/130/776 Wilman R.: J/A+A/471/1105 J/ApJ/591/640 Wilms J.: J/A+A/342/736 J/A+A/395/129 J/A+A/407/1039 J/A+A/443/103 J/A+A/447/245 Wilmsen U.: J/A+AS/70/481 Wils P.: II/242 J/MNRAS/368/1757 J/other/IBVS/5586 Wilson A.S.: J/A+A/409/115 J/A+A/435/521 J/ApJ/591/138 J/ApJ/601/735 J/ApJS/105/93 Wilson C.A.: J/ApJS/154/585 WILSON C.D.: J/AJ/105/499 Wilson C.D.: J/AJ/110/2715 J/ApJ/599/1049 J/ApJ/625/891 J/ApJ/644/990 Wilson D.M.: J/A+A/467/785 J/MNRAS/372/1117 J/MNRAS/379/647 J/MNRAS/379/816 Wilson G.: J/A+A/371/833 J/A+A/398/901 J/AJ/131/250 J/ApJS/154/155 Wilson J.: J/AJ/119/486 Wilson J.C.: J/ApJ/635/280 Wilson J.W.: J/ApJ/525/420 J/PASP/105/36 Wilson N.J.: J/A+A/432/731 Wilson R.: II/59B II/86 WILSON R.: III/39A V/38 Wilson R.E.: J/A+A/452/959 J/AJ/110/1350 J/ApJ/484/394 Wilson R.F.: III/21 Wilson R.W.: J/ApJ/517/282 VIII/10 VIII/28 Wilson T.: J/ApJS/107/175 Wilson T.L.: J/A+A/314/615 J/A+A/325/255 J/A+A/327/1177 J/A+A/441/181 Wilson W.E.: J/MNRAS/354/305 J/MNRAS/371/898 Wilson W.J.F.: J/PASP/106/1120

Windhorst R.A.: J/A+A/397/361 J/A+AS/58/1 J/A+AS/58/39 J/A+AS/73/103 J/AJ/112/628

J/AJ/117/2757 J/AJ/123/2402 J/AJ/126/1286 J/AJ/129/26 WINDHORST R.A.: J/ApJ/405/498

Windhorst R.A.: J/ApJ/453/48 J/ApJ/612/L93 J/ApJ/630/784 J/ApJ/639/724 J/MNRAS/317/801

J/MNRAS/358/1159 Wing R.F.: J/AJ/119/1424 Wingert D.W.: J/ApJ/639/1069 Winget D.E.: J/AJ/128/1825 J/AJ/131/582 Winkler C.: J/A+AS/143/145 J/ApJ/636/765 J/ApJS/170/175 WINKLER H.: J/A+AS/93/383 Winkler H.: J/other/SAAOC/15.78 Winkler L.I.: J/AJ/121/2148 WINKLER P.F.: J/AJ/107/1453 Winn J.N.: J/AJ/129/2856 J/AJ/130/1977 J/ApJS/145/83 Winnberg A.: J/A+AS/128/35 J/A+AS/92/43 WINNBERG A.: J/A+AS/93/121 Winnberg A.: J/AZh/79/610 Winnewisser G.: J/A+A/309/581 J/A+A/425/767 Winnewisser G: J/A+AS/144/123 Winnick R.: J/AJ/127/840 Winstanley N.: J/A+A/472/805 Winston E.: J/A+A/463/275 Winter E.L.: J/ApJ/616/1284 Winter L.: I/268 I/273A I/289 Winters J.M.: J/A+A/361/641 J/A+A/376/997 J/A+A/475/559 WIRAMIHARDJA S.D.: III/177 III/177 III/177 III/177 Wiren S.: J/A+A/427/769 J/A+A/440/409 J/A+AS/132/305 Wirstrom E.S.: J/A+A/473/177 Wirtanen A.C.: I/11 Wirth G.D.: J/AJ/127/3121 J/ApJ/565/208 J/ApJS/129/475 Wischnewski R.: J/other/APh/26.282 Wischnjewsky M.: J/AJ/112/2408 J/ApJS/133/119 Wise J.H.: J/ApJ/613/129 Wise M.W.: J/ApJ/405/94 WISHNJEWSKY M.: J/AJ/106/2575 Wisniewski J.P.: J/A+A/408/305 J/ApJ/598/L43 Wisniewskj W.Z.: II/5A

Wisotzki L.: II/253 J/A+A/299/39 J/A+A/358/77 J/A+A/363/1040 J/A+A/375/366 J/A+A/377/442

J/A+A/431/143 J/A+A/440/53 J/A+A/466/41 WISOTZKI L.: J/A+AS/110/469 Wisotzki L.: J/A+AS/115/235 J/A+AS/128/507 J/MNRAS/313/641 Wissing H.: J/other/APh/26.282 with assist. of Manzolini F.: II/61 Withford A.E.: II/8 Witt A.N.: J/A+AS/127/1 Witteborn F.C.: J/AJ/110/275 J/AJ/112/2274 J/AJ/112/241 Wittek W.: J/A+A/390/39 J/A+A/406/L9 Witthoeft M.C.: J/A+A/466/763 Wittkowski M.: J/A+A/466/1099 J/A+A/470/191 WITTMANN A.D.: J/A+AS/103/293 WITZEL A.: I/169 Witzel A.: J/A+A/402/151 J/A+A/476/L17 J/A+A/85/329 J/ApJ/439/983 WITZEL A.: VII/78 VIII/5 Witzel: J/A+A/364/391 Wlasuk P.: V/36B Woitas J.: J/A+A/369/249 J/A+A/376/982 Wold M.: J/MNRAS/375/989 Wold T.: J/MNRAS/373/781 WOLF B.: J/A+A/289/740 Wolf B.: J/A+A/389/931 J/A+AS/133/299 Wolf C.: II/253 J/A+A/377/442 J/ApJS/155/73 Wolf G.W.: J/A+A/459/557

Wolf M.: J/A+A/387/580 J/A+A/455/1037 J/A+A/463/1061 J/A+A/463/233 J/A+A/464/263

J/A+A/465/331 Wolf S.: J/AJ/129/1049 J/PASP/118/1690 Wolfe A.M.: J/AJ/123/2206 J/ApJ/543/552 J/ApJ/635/123 J/PASP/118/1077 Wolfe C.: VIII/42 Wolfe M.A.: J/AJ/131/1674 Wolff M.J.: J/ApJ/635/560 Wolff S.C.: J/AJ/129/809 J/ApJ/601/979 J/PASP/109/759 Wolfgang Hovest: VII/199 Wolfire M.G.: J/ApJ/635/560 Wolk S.: J/ApJS/160/319 Wolk S.J.: J/A+AS/146/323 J/AJ/128/2316 J/AJ/131/527 J/ApJ/582/382 J/ApJ/588/1009 J/ApJ/589/509

J/ApJ/606/466 Wollack E.: J/ApJS/148/97 Wolleben M.: J/A+A/448/411 WOLSTENCROFT R.D.: VII/169 Wolter A.: IX/15 J/A+A/435/799 J/ApJ/634/272 J/ApJS/99/701 Womble D.S.: J/ApJ/472/509 Wong O.I.: J/AJ/124/690 Wong S.K.: J/AJ/111/846 Wong T.: J/AJ/116/2341 J/MNRAS/367/1609 Wonnacott D.: J/MNRAS/274/1194 J/MNRAS/311/456 Wood D.O.S.: J/ApJS/91/659 J/ApJS/95/457 WOOD F.B.: VI/23 Wood F.B.: VI/44 Wood K.: J/AJ/131/1574 Wood K.M.: J/ApJ/449/164 WOOD K.S.: J/ApJS/56/403 VII/80 Wood P.R.: J/A+AS/134/489 J/A+AS/146/217 J/AJ/108/932 J/AJ/123/840 J/AJ/99/784 WOOD P.R.: J/ApJS/92/125 Wood P.R.: J/MNRAS/376/313 J/PASJ/53/293 J/PASJ/54/719 Wooden D.H.: J/AJ/110/275 J/AJ/112/2274 J/AJ/112/241 Woodgate B.: J/ApJ/612/481 Woodgate B.E.: J/ApJ/548/127 J/ApJ/614/75 J/ApJS/146/209 Woodroofe M.: J/AJ/131/2114 Woods D.F.: J/AJ/132/197 Woods P.M.: J/ApJ/615/887 Woods T.: V/68A Wood-Vasey W.M.: J/AJ/130/2453 J/AJ/131/1648 Woodward C.E.: J/AJ/107/2120 Woolf V.M.: J/ApJ/617/1091 Woolley R.: V/32A V/4 WOOLSEY E.G.: I/20 Woon Kang Y.: J/A+A/430/255 Woosley S.: J/other/GCN/1070.1 WORKING GROUP OF CGSC: I/179 Working Group of GCPA: I/180 WORLEY C.E.: I/107A Worley C.E.: I/237 II/235 J/AJ/116/917 J/AJ/117/1890 J/AJ/117/1905 J/AJ/117/2511 J/AJ/118/1395

V/39 WORLEY Ch.: VI/18 Worley: V/111 Worrall D.: J/A+A/476/1191 Worrall D.M.: J/ApJS/92/53 Worthey G.: J/ApJS/111/377 J/ApJS/116/1 J/ApJS/160/176 WORTHEY G.: J/ApJS/94/687 Woschnagg K.: J/other/APh/26.282 Woudt P.A.: J/A+A/352/39 J/A+A/380/441 J/A+A/415/9 J/A+A/466/481 J/A+AS/127/463 J/MNRAS/354/980

J/MNRAS/366/267 J/MNRAS/368/534 J/MNRAS/369/1131 Wouterloot J.G.A.: J/A+A/309/581 J/A+A/430/549 J/A+A/464/909 J/A+AS/144/123 J/A+AS/80/149

J/A+AS/98/589 Wozniak H.: J/A+A/465/1099

Wozniak P.: J/AcA/48/147 J/AcA/48/175 J/AcA/48/563 J/AcA/49/223 J/AcA/49/319 J/AcA/49/437

J/AcA/49/521 J/AcA/50/1 J/AcA/52/217 J/AcA/53/93 J/MNRAS/348/1439 Wozniak P.R.: J/AcA/50/421 J/AcA/51/175 J/AcA/51/317 J/AcA/52/129 J/AcA/52/143 J/AcA/53/1

J/AJ/127/2436 J/ApJ/636/240 J/MNRAS/327/601 J/MNRAS/340/1346 J/MNRAS/356/331 J/MNRAS/370/435 Wozniak S.J.: J/AJ/128/2965 Wray J.D.: III/86 Wray J.J.: J/MNRAS/349/1059 Wren J.: J/AJ/119/1901 J/AJ/127/2436 Wright A.: VIII/15 Wright A.E.: J/AJ/123/873 J/AJ/124/1954 J/AJ/124/690 J/AJ/128/16 WRIGHT A.E.: J/MNRAS/251/330 Wright A.E.: J/MNRAS/342/738 J/MNRAS/346/787 J/MNRAS/355/1303 WRIGHT A.E.: VIII/27 VIII/27 VIII/27 Wright A.E.: VIII/38 VIII/38 VIII/38 VIII/38 VIII/56 Wright C.M.: J/A+A/450/253 Wright C.O.: III/231 V/114 Wright E.L.: J/ApJS/148/97 Wright G.A.: J/ApJ/517/282 Wright J.T.: J/ApJS/152/261 Wrobel J.M.: J/AJ/130/923 J/ApJS/145/83 VIII/60

Wroblewski H.: J/A+A/18/341 J/A+A/367/725 J/A+AS/105/179 J/A+AS/110/27 J/A+AS/115/481 J/A+AS/128/457 J/A+AS/139/25 J/A+AS/142/369 J/A+AS/78/231 J/A+AS/83/317 J/A+AS/91/129 J/A+AS/92/449 J/other/POAN/2.59

Wroblewski W.: J/A+AS/122/447 Wu C.: J/A+A/466/627 J/AJ/130/1640 Wu C.-C.: J/ApJS/82/197 Wu H.: J/A+A/397/361 J/A+AS/127/521 J/A+AS/132/181 J/AJ/112/628 J/AJ/117/2757 J/AJ/122/1718

J/AJ/126/1286 J/AJ/127/2579 J/ApJ/510/659 J/ApJ/625/78 J/ApJS/115/185 Wu J.: J/A+A/380/665 J/A+A/449/143 J/AJ/129/1818 Wu J.-H.: J/AJ/123/578 Wu J.H.: J/MNRAS/373/209 Wu J.-H.: J/other/PPMtO/17b.1 Wu K.: J/A+A/410/53 J/A+A/459/777 J/ApJS/144/213 J/ApJS/154/519 J/ApJS/159/41 Wu K.L.: J/AJ/123/2223 J/ApJS/157/175 Wu P.: J/A+A/474/783 Wu W.: J/ApJS/143/377 Wu X.: II/203A II/203A IX/35 IX/35 J/ApJS/93/569 J/ApJS/96/461 J/MNRAS/348/1439 Wu X.-B.: J/A+A/466/627 J/ApJ/614/91 J/ApJ/645/890 Wu X.J.: J/MNRAS/328/855 J/MNRAS/362/1189 Wu X.-P.: J/A+A/338/813 Wu Y.: J/A+A/380/665 J/A+A/426/503 J/A+A/461/197 J/A+AS/115/283 Wu Y.F.: J/A+AS/127/181 Wu Y.-F.: J/other/PPMtO/14.185 Wu Z.: J/A+A/449/143 J/A+A/466/63 Wu Z.Y.: J/A+A/381/464 Wu Z.-Y.: J/A+AS/142/373 J/PASP/118/1104 Wucknitz O.: J/A+A/401/927 Wuilleumier F.J.: J/A+A/439/387 W.Wittek.: J/A+A/410/813 Wyckoff E.: J/ApJS/122/81 Wycoff G: I/259 Wycoff G.L.: I/222 I/246 I/247 I/268 I/275 I/276 I/289 I/294 I/297 II/235 IV/25 J/AJ/117/1890

J/AJ/120/1120 J/AJ/122/1586 J/AJ/123/1023 J/AJ/124/2254 J/AJ/127/539 J/AJ/128/3012 J/AJ/132/50 Wyder T.K.: J/A+A/357/437 J/AJ/130/825 J/ApJ/619/L95 J/ApJS/164/38 J/PASP/111/685 Wyk F.V.: J/MNRAS/325/931 Wyrowski F.: J/A+A/466/1065 J/A+A/475/925

Wyrzykowski L.: II/278 J/AcA/51/221 J/AcA/51/317 J/AcA/52/143 J/AcA/52/217 J/AcA/53/1 J/AcA/53/93 J/AcA/54/347 J/ApJ/636/240 J/MNRAS/348/1439 J/MNRAS/356/331 J/MNRAS/370/435

WYSE R. F.: III/167 Wyse R.F.G.: J/AJ/109/1095 J/AJ/113/634 J/AJ/131/895

Xi X.J.: J/MNRAS/376/1707 Xia X.: J/AJ/117/2757 J/MNRAS/308/897 Xia X.Y.: J/A+AS/127/521 J/A+AS/132/181 J/ApJ/625/78 Xiang D.-L.: J/other/PPMtO/14.185 Xiang F.Y.: J/A+AS/124/281 Xie G. Z.: J/A+A/278/6 XIE G.Z.: J/A+AS/106/361 Xie G.Z.: J/A+AS/132/83 J/A+AS/136/455 J/ApJ/522/846 Xilouris K.M.: J/MNRAS/359/1524 Xiong Y.: J/ApJ/525/420 XMM Science Team: B/xmm XMM-Newton Survey Science Centre Consortium: IX/37 IX/39 Xu C.: J/PASP/113/10 Xu C.K.: J/AJ/129/1198 J/AJ/129/2074 J/ApJ/619/L95 Xu D.-W.: J/A+A/414/487 Xu D.W.: J/A+AS/139/575 J/ApJ/590/73 Xu J.: J/A+AS/132/233 Xu K.: J/ApJS/164/38 Xu W.: J/A+A/338/813 J/A+A/472/763 J/AJ/117/2757 J/ApJS/98/1 J/ApJS/98/33 J/ApJS/99/297 Xu X.W.: J/other/APh/26.282 Xu Y.: J/ApJS/166/526 J/other/ChJAA/3.49 Xu ZHOU: II/262 Xue S.: J/A+A/397/361 J/AJ/117/2757 J/AJ/122/1718 J/AJ/126/1286 J/AJ/127/2579 Xue S.-J.: J/ApJ/510/659

Yadav R.K.S.: J/MNRAS/337/133 Yagi H.: J/AJ/123/1528 Yagi M.: J/ApJS/137/279 J/MNRAS/357/1357 J/PASJ/55/1079 YAGUDIN L.I.: I/159A Yagudin L.I.: I/195 J/PAZh/32/628 Yahata N.: J/AJ/123/2206 J/ApJ/538/493 J/ApJS/135/41 YAHIL A.: II/174 Yahil A.: J/ApJ/513/34 J/ApJ/538/493 VII/185 VII/185 VII/185 YAKOVLEV D.G.: J/A+AS/109/125 Yakovleva V.A.: J/AZh/71/717 Yakut K.: J/A+A/405/1087 J/A+A/463/243 J/ApJ/629/1055 Yamada K.M.T.: J/A+A/425/767 Yamada R.: J/MNRAS/330/137 Yamada S.F.: J/PASJ/57/881 YAMADA T.: J/A+A/299/347 Yamada T.: J/A+AS/104/529 J/AJ/111/42 J/ApJ/518/656 J/ApJ/528/123 J/ApJ/634/861 J/ApJS/107/521

J/ApJS/112/245 J/ApJS/148/275 J/ApJS/89/57 J/MNRAS/307/236 J/MNRAS/379/1343 J/PASJ/57/165 Yamagata T.: J/ApJS/106/463 Yamaguchi N.: J/PASJ/53/971 Yamaguchi R.: J/PASJ/51/791 J/PASJ/53/971 Yamamura I.: J/A+A/376/997 J/ApJS/106/463 J/PASP/117/199 Yamamura Y.: J/A+A/403/943 Yamashita A.: J/PASJ/53/445 Yamashita N.: J/PASJ/52/631 Yamashita Y.: II/192 Yamauchi S.: J/ApJ/635/214 J/ApJS/134/77 Yamazaki R.: J/ApJ/609/935 Yan H.: J/AJ/117/2757 J/ApJ/612/L93 Yan H.J.: J/AJ/112/628 Yan L.: J/AJ/112/1489 J/AJ/130/1324 J/AJ/131/250 J/ApJ/586/765 Yan R.: J/ApJ/625/6 YAN Y.: VI/80 Yan Y.-H.: J/other/AcApS/16.103 Yan Y.R.: J/A+A/454/379 J/MNRAS/376/1707 Yanagisawa K.: J/PASJ/57/13 Yanagisawa T.: J/MNRAS/330/137 J/MNRAS/331/609 Yang H.T.: J/other/NewA/10.325 Yang J.: J/AJ/117/446 J/ApJ/596/1064 J/ApJ/629/288 J/ApJS/141/157 Yang S.: J/A+A/312/879 J/A+A/319/867 J/A+A/367/236 J/A+A/387/580 J/A+A/408/611 J/A+A/416/669

J/A+A/446/583 J/A+A/455/1037 J/A+A/459/849 J/A+A/463/1061 J/A+A/463/233 Yang S.L.S.: III/185 Yang T.-G.: I/277 I/288 Yang Y.: J/A+A/449/143 J/A+AS/125/475 J/A+AS/136/139 J/AJ/128/1483 J/AJ/128/1501 Yang Y.-G.: J/AJ/130/2252 Yanny B.: J/AJ/107/1745 YANNY B.: J/AJ/108/1786

Yanny B.: J/AJ/113/669 J/AJ/116/1757 J/AJ/121/2308 J/AJ/121/2331 J/AJ/122/503 J/AJ/123/567 J/AJ/126/1023 J/AJ/126/2209 J/AJ/126/2579 J/AJ/127/1811 J/AJ/130/873 J/ApJ/409/28 J/ApJ/588/824 J/ApJS/145/245

Yao Y.: J/AJ/117/446 J/ApJ/596/1064 J/ApJ/629/288 Yao Y.-Q.: J/other/ChA+A/28.308 Yasaka Y.: J/PASJ/57/13 Yasuda H.: I/249 Yasuda N.: J/ApJS/137/279 J/MNRAS/357/1357 J/PASJ/55/1079 J/PASJ/57/165 Yates F.F.: II/96 Yates J.A.: J/MNRAS/273/411 J/MNRAS/282/779 Yatsenko A.: I/261 Yatsenko A.I.: I/285 J/A+A/384/145 J/A+AS/121/243 Yatsevich A.V.: J/AZh/72/495 YATSKIV Y.S.: I/159A Yearsley J.M.: J/MNRAS/279/595 Yee H.K.C.: J/ApJ/471/694 J/ApJ/565/208 J/ApJ/582/1123 J/ApJS/102/289 J/ApJS/113/1 J/ApJS/116/211

J/ApJS/116/231 J/ApJS/116/247 J/ApJS/129/475 J/ApJS/157/1 J/ApJS/157/228 J/ApJS/158/161 Yentis D.: J/ApJS/140/239 Yentis D.J.: I/277 I/288 YENTIS D.J.: VII/80 Yerbury M.J.: VIII/61 Yerli S.K.: J/PAZh/26/843 Yi S.: J/ApJ/633/871 Yi S.K.: J/ApJS/162/1 Yim H.S.: J/A+A/450/39 Yin Q.F.: J/AJ/125/2411 J/ApJ/378/65 J/ApJS/159/282 VIII/65 Yin S.Y.: J/A+A/462/535 Yip C.-W.: J/AJ/131/84 Yngvason J.: J/ApJ/416/276 Yock P.C.M.: J/MNRAS/330/137 J/MNRAS/331/609 Yodh G.: J/other/APh/26.282 Yokogawa J.: J/PASJ/55/161 Yokoyama K.: J/PASJ/52/631 Yonekura Y.: J/ApJS/110/21 J/ApJS/95/419 J/PASJ/53/971 J/PASJ/56/313 Yonetoku D.: J/ApJ/609/935 Yong D.: J/AJ/130/1111 J/AJ/131/2274 J/ApJ/638/1018 J/PASP/115/22 Yoon S.-J.: J/AJ/122/3219 York D. G.: J/ApJS/122/355 York D.G.: II/149 III/157 J/A+A/423/449 J/AJ/121/2308 J/AJ/121/2331 J/AJ/121/31 J/AJ/122/503

J/AJ/123/3409 J/AJ/123/567 J/AJ/124/1651 J/AJ/125/1711 J/AJ/125/1817 J/AJ/126/1023 J/AJ/126/2209 J/AJ/126/2579 J/AJ/127/2838 J/AJ/127/899 J/AJ/129/2542 J/AJ/131/2018 J/ApJ/588/824 J/ApJ/607/426 J/ApJS/146/1 J/ApJS/155/257

Yoshida A.: J/other/GCN/1070.1 Yoshida M.: J/ApJ/634/861 J/PASJ/57/165 Yoshida S.: II/273 YOSHIDA S.: III/177 III/177 III/177 III/177 Yoshida S.: J/other/APh/26.282 Yoshii Y.: J/AJ/115/168 J/AJ/119/2866 J/ApJ/475/445 J/ApJS/141/23 J/ApJS/141/31 Yoshimura A.: J/PASJ/52/631 Yoshino K.: VI/61 YOSHIZAWA M.: I/167 I/186 I/187 I/188 I/198 Yoshizawa M.: I/248 J/MNRAS/330/137 Yoss K.M.: J/AJ/122/248 J/PASP/111/1169 V/103 V/33 YOSS K.M.: V/7 Yost S.A.: J/ApJS/165/19 You J.: J/A+A/362/223 You X.-P.: J/other/ChJAA/6.237 Youn J.H.: J/A+A/404/621 Youn J.-H.: J/A+A/405/231 Young C.K.: J/A+AS/127/367 J/MNRAS/279/595 Young E.: J/AJ/131/1574 J/ApJ/566/993 J/ApJ/638/897 J/PASP/118/1690 Young E.T.: II/126 J/AJ/126/1665 YOUNG E.T.: J/ApJ/397/520 Young E.T.: J/ApJ/540/1016 J/ApJ/620/1010 Young K.E.: J/ApJ/638/293 Young K.J.: J/A+A/429/317 Young P.J.: J/AJ/81/807 Young P.R.: J/A+AS/125/149 Younger F.: J/AJ/84/1305 YOUNGER P.F.: II/133 Younger P.F.: J/AJ/130/1194 Yu L.H.: J/AJ/130/23 Yu W.: J/A+A/426/503 Yu Y.: J/A+A/454/379 Yuan Q.: J/AJ/122/1718 J/ApJS/149/53 Yuan Q.-R.: J/other/PPMtO/17b.1 Yuan Tang: VIII/69A Yuan W.: J/A+A/319/413 J/A+A/330/108 J/A+A/476/1191 Yuan W.-M.: J/other/PPMtO/17b.1 Yuce K.: J/A+A/371/1078 J/BaltA/14/51 Yudin B.F.: J/A+A/311/484 J/A+A/323/113 J/A+A/394/617 YUDIN B.F.: J/A+AS/93/383 Yudin B.F.: J/AZh/71/618 Yudin R.V.: J/A+A/368/912 J/A+A/384/504 J/A+A/412/405 J/A+AS/131/401 Yuk I.-S.: J/AJ/126/2840 Yun J.L.: J/A+A/467/1397 J/A+A/470/231 J/A+A/471/573 Yun Min S.: J/ApJ/554/803 Yun Min Su: J/ApJS/63/821 Yungelson L.: J/A+A/378/556 YUNGELSON L.: V/41 Yungelson L.R.: J/AZh/79/738 Yurchenko A.V.: J/AZh/71/717 Yusef-Zadeh F.: J/ApJS/155/421 Yushchenko A.: J/A+A/430/255 J/MNRAS/359/865 Yushchenko A.V.: J/PASP/118/636 Yushchenko V.: J/A+A/430/255 Yushkin M.V.: J/AZh/82/601 Yusup A.: J/MNRAS/328/855 Yutani M.: II/192

Zabludoff A.: J/AJ/105/788 J/MNRAS/360/587 Zabludoff A.I.: J/AJ/106/1273 J/ApJ/496/39 J/ApJ/539/136 J/ApJ/549/172 J/ApJ/591/764 J/ApJ/619/193

J/MNRAS/327/588 J/MNRAS/366/645 Zabolotskikh M.V.: J/PAZh/23/90 J/PAZh/25/689 Zacarias N.: II/246 Zaccardo A.: J/A+A/417/499 Zacharias M.I.: I/265 I/268 I/289 J/AJ/125/2728 Zacharias N.: I/265 I/268 I/273A I/289 I/294 I/297 J/AJ/110/880 J/AJ/125/2728 Zacs L.: J/A+AS/136/453 Zagaria M.: J/A+A/385/39 ZAGGIA S.: J/A+A/296/319 Zaggia S.: J/A+A/333/479 J/A+A/341/641 J/A+A/379/436 J/A+A/452/119 J/A+A/456/881 J/A+A/462/553

J/A+A/464/201 J/A+A/465/815 ZAGGIA S.: J/A+AS/105/433 Zaggia S.: J/A+AS/133/325 J/A+AS/137/75 J/A+AS/137/83 ZAIME D.: II/198 Zaitseva G.V.: J/PAZh/23/277 J/PAZh/26/13 J/PAZh/31/116 Zajac J.M.: J/AJ/131/527 Zajtseva G.V.: J/A+A/341/553 Zakamska N.L.: J/AJ/126/1720 J/AJ/126/2125 J/AJ/128/1002 J/AJ/129/2542 Zakharova L.A.: J/AZh/71/588 Zakhozhaj V.A.: J/other/KFNT/12.20 V/101 Zakirov M.: J/A+A/343/894 J/A+A/374/638

Zakirov M.M.: J/PAZh/23/458 J/PAZh/23/626 J/PAZh/24/145 J/PAZh/25/276 J/PAZh/26/525

J/PAZh/27/442 J/PAZh/27/632 Zaldarriaga M.: J/AJ/129/2856 Zalles R.: J/PAZh/32/628 Zalles R.F.: I/195 Zalubas R.: VI/74 VI/74 Zalutsky V.T.: J/A+AS/132/233 Zamanov R.: J/A+AS/114/499 J/ApJS/145/199 Zambelli G.: J/MNRAS/324/1131 Zambon M.: VII/3 Zamkoff E.: J/AJ/116/180 Zamorani G.: J/A+A/329/482 J/A+A/342/1 J/A+A/371/833 J/A+A/379/35 J/A+A/392/395 J/A+A/398/901

J/A+A/403/857 J/A+A/428/1043 J/A+A/432/69 J/A+A/437/883 J/A+A/457/501 J/A+A/457/79 J/A+A/466/41 J/A+AS/130/323 J/AJ/127/3075 J/ApJS/121/473 J/MNRAS/296/599 J/MNRAS/312/540

Zamorano J.: J/A+A/365/370 J/A+A/379/798 J/A+AS/118/7 J/A+AS/120/323 J/A+AS/141/409

J/AJ/105/427 J/ApJS/105/343 J/ApJS/122/415 J/ApJS/95/387 J/MNRAS/338/508 J/MNRAS/338/525 Zanichelli A.: J/A+A/379/21 J/A+A/379/35 J/A+A/403/857 J/A+A/428/1043 J/A+A/457/79

J/A+A/463/519 J/ApJ/591/640 Zanin C.: J/A+AS/137/293 Zapatero Osorio M.R.: J/A+A/409/169 J/A+A/419/167 J/AN/325/705 Zappacosta L.: J/A+A/434/801 J/A+A/468/807 Zappala V.: VII/108C Zaritsky D.: J/A+A/427/397 J/A+A/444/365 J/AJ/121/891 J/AJ/123/855 J/AJ/127/1531 J/AJ/128/1606

J/AJ/129/2701 J/AJ/131/414 J/ApJ/563/629 J/ApJ/566/103 J/ApJ/604/141 J/ApJ/630/206 J/ApJS/137/117 J/MNRAS/371/1777 Zarka P.: J/A+A/475/359 Zaroubi S.: J/MNRAS/352/939 Zasov A.: J/A+AS/127/117 Zasov A.V.: J/AZh/78/963 Zaum A.: J/A+AS/121/247 J/MNRAS/306/279 ZAVAGNO A.: J/A+A/303/175 Zavagno A.: J/A+A/344/943 J/A+A/360/1107 J/A+A/394/225 J/A+A/399/1135 Zavala J.: J/A+A/412/633 Zavala R.T.: J/ApJ/589/733 Zavlin V.E.: J/ApJ/578/405 Zboril M.: J/A+A/324/949 J/A+A/363/1051 J/A+A/373/987 J/A+A/406/193 Zdanavicius J.: J/BaltA/10/349 J/BaltA/11/441 J/BaltA/11/75 J/BaltA/13/555 J/BaltA/14/1 Zdanavicius K.: II/93 J/BaltA/10/349 J/BaltA/11/441 J/BaltA/11/75 J/BaltA/13/555 J/BaltA/14/1 Zealey W.: J/MNRAS/339/105 Zebrun K.: II/278 J/AcA/48/147 J/AcA/48/175 J/AcA/48/289 J/AcA/48/563 J/AcA/49/223 J/AcA/49/319

J/AcA/49/437 J/AcA/49/521 J/AcA/50/1 J/AcA/51/175 J/AcA/51/221 J/AcA/51/317 J/AcA/52/129 J/AcA/52/143 J/AcA/52/217 J/AcA/53/1 J/AcA/53/93 J/AcA/54/347 J/ApJ/636/240 J/MNRAS/348/1439 J/MNRAS/356/331

ZEILIK M.: V/71

Zeilinger W.W.: J/A+A/374/394 J/A+A/433/497 J/A+A/445/79 J/A+A/463/455 J/A+AS/120/257

J/A+AS/131/105 J/AJ/129/1832 J/MNRAS/364/146 Zeippen C.J.: J/A+A/372/1078 J/A+A/372/1083 J/A+A/394/753 ZEIPPEN C.J.: J/A+AS/102/435

Zeippen C.J.: J/A+AS/108/1 J/A+AS/120/361 J/A+AS/123/159 J/A+AS/131/499 J/A+AS/135/159

J/A+AS/143/483 J/A+AS/144/141 Zeippen: J/A+AS/137/165 Zekser K.: J/ApJ/621/53 Zemcov M.: J/MNRAS/376/1073 Zeng F.: J/A+AS/118/1 J/A+AS/136/1 Zeng F.M.: J/A+A/430/327 Zeng Q.: J/ApJ/629/288 Zeng Z.: J/A+AS/118/1 J/A+AS/136/1 Zenina I.A.: J/A+A/323/775 Zensus J.A.: J/A+A/476/L17 J/AJ/130/2473 J/ApJ/609/539 Zensus: J/A+A/364/391 Zentner A.R.: J/ApJ/645/228 Zepf S.E.: J/A+A/448/155 J/AJ/118/752 J/ApJ/513/733 J/ApJ/586/814 J/ApJS/110/1 J/MNRAS/355/608 Zerbi F.M.: J/A+A/336/920 J/A+A/469/503 J/A+A/474/793 Zezas A.: J/ApJS/161/271 Zhai D.S.: J/A+AS/122/515 Zhang B.: J/A+A/397/257 J/A+A/462/535 J/other/PBeiO/16.18 Zhang C.Y.: J/ApJS/98/659 Zhang F.: J/MNRAS/370/1181 Zhang H.: J/AJ/127/2579 Zhang H.L.: J/A+A/283/319 J/A+A/293/953 J/A+AS/119/523 J/A+AS/123/575 J/A+AS/126/373 Zhang H.W.: J/A+A/367/1049 J/A+A/425/671 J/A+A/449/127 J/A+AS/141/491 Zhang J.: J/MNRAS/328/855 Zhang J.S.: J/A+A/407/899 Zhang P.: J/AJ/131/1942 J/other/NewA/10.325 J/other/NewA/9.509 Zhang Q.: J/A+A/461/197 J/A+A/466/1065 J/A+A/466/989 J/A+A/475/925 J/AJ/124/1418 ZHANG Q.: J/ApJ/416/787 Zhang Q.: J/ApJ/561/727 J/ApJ/606/929 Zhang R.X.: J/A+A/395/587 J/A+AS/122/515 Zhang T.: J/A+A/449/143 J/AJ/131/2245 Zhang X.: J/A+A/395/587 J/A+AS/132/83 J/A+AS/136/455 J/A+AS/99/545 J/ApJ/522/846 VIII/44 Zhang X.B.: J/A+A/455/247 Zhang X.-Z.: J/other/AcApS/16.103 Zhang Y.: J/A+A/404/545 J/A+A/442/249 J/MNRAS/368/1959 Zhang Y.H.: J/A+A/278/6 J/A+A/381/757 ZHANG Y.H.: J/A+AS/106/361 Zhang Z.: J/A+AS/132/233 Zhang Z.-S. Li Y.-I.: J/A+AS/131/115 Zhao C.: J/ApJS/145/245 Zhao G.: J/A+A/305/245 J/A+A/367/1049 J/A+A/397/257 J/A+A/423/683 J/A+A/425/671 J/A+A/430/327

J/A+A/449/127 J/A+AS/118/1 J/A+AS/141/491 J/ApJS/126/461 J/AZh/83/265 J/MNRAS/370/2091 J/MNRAS/373/209 J/PASJ/58/389

Zhao H.: J/MNRAS/339/701 Zhao J.: I/258 Zhao J.H.: I/257B Zhao J.-H.: J/AJ/127/3399 J/AJ/98/64 ZHAO J.L.: I/215 Zhao J.L.: J/A+A/381/464 J/A+A/426/819 J/A+AS/100/243 J/A+AS/105/15 J/A+AS/118/503 Zhao J.-L.: J/A+AS/128/255 Zhao J.L.: J/A+AS/131/239 Zhao J.-L.: J/A+AS/131/89 Zhao J.L.: J/A+AS/133/387 Zhao J.-L.: J/other/ChJAA/6.435 Zhao M.: J/A+A/426/503 Zhao P.: J/AJ/131/527 Zhao X.: J/ApJ/626/L5 Zhao Y.H.: J/AJ/130/1640 Zhaoji JIANG: II/262 Zharova A.V.: J/PAZh/28/201 Zhekanis G.V.: J/A+A/370/78 J/A+AS/139/545 Zhelenkova O.P.: J/AZh/72/613 J/AZh/79/589 J/other/BSAO/40.5 Zheleznyak A.P.: J/A+A/420/447 Zheleznyakov V.V.: J/PAZh/30/430 Zhen Y.: J/A+AS/99/545 Zheng W.: J/A+A/451/457 J/AJ/121/2308 J/AJ/127/3553 J/AJ/128/1990 J/AJ/131/2722 J/ApJ/475/469

J/ApJ/612/202 J/ApJ/621/53 J/ApJS/115/185 J/ApJS/139/369 J/ApJS/150/1 J/ApJS/155/271 J/ApJS/155/73 Zheng X.-W.: J/ApJS/160/220 Zheng X.W.: J/ApJS/166/526 Zheng X.-W.: J/other/ChJAA/3.49 Zheng X.Z.: J/A+A/430/115 J/A+A/435/507 J/A+A/465/1099 J/ApJ/590/73 Zheng Y.: VIII/44 Zheng Y.-J.: J/other/AcApS/16.103 Zheng Z.: J/AJ/117/2757 J/AJ/117/2757 Zheng Z.Y.: J/AJ/112/628 Zhili L.: J/A+AS/114/337 Zhou A.Y.: J/A+A/374/1056 Zhou A.-Y.: J/A+A/456/261 Zhou H.-Y.: J/ApJ/645/856 J/other/ChJAA/2.501 Zhou X.: J/A+A/397/361 J/A+A/449/143 J/AJ/112/628 J/AJ/117/2757 J/AJ/122/1718 J/AJ/126/1286

J/AJ/127/2579 J/AJ/129/1818 J/AJ/131/2245 J/ApJ/645/488 J/ApJS/149/53 J/MNRAS/373/209 J/PASP/118/1104 Zhou X.-L.: J/ApJ/618/L83 Zhu J.: J/A+A/397/361 J/AJ/112/628 J/AJ/117/2757 J/AJ/122/1718 J/AJ/126/1286 J/other/AcApS/16.103 Zhu L.-Y.: J/AJ/130/2252 J/AJ/131/1032 Zhuchkov R.Ya.: J/AZh/82/231 Ziaeepour H.: J/A+A/476/1191 Zibetti S.: J/A+A/430/411 Zickgraf F.-J.: J/A+A/406/535 J/A+A/414/487 J/A+A/433/151 J/A+A/439/129 J/A+AS/132/341

J/AJ/132/137 J/ApJS/117/319 Zieba S.: J/A+A/303/420 Ziegler B.: J/A+A/398/49 J/A+A/421/41 J/ApJS/117/319

Ziegler B.L.: J/A+A/418/885 J/A+A/420/97 J/A+A/421/59 J/A+A/422/907 J/MNRAS/337/256

J/MNRAS/358/233 Ziegler M.: J/A+A/470/317 Ziener R.: J/AJ/107/2101 Zigman V.: J/A+A/477/967 Zijlstra A.: J/A+AS/132/13 Zijlstra A.A.: J/A+A/406/51 ZIJLSTRA A.A.: J/A+AS/108/485 Zijlstra A.A.: J/A+AS/125/289 J/A+AS/125/419 J/MNRAS/376/313 Zimmerman N.W.: I/45 Zimmermann H.-U.: IX/10A J/A+A/378/30 J/other/SerAJ/166.13Short Zimmermann M.: J/A+AS/109/201 Zimmermann U.: IX/11 Zimmermann U. Max-Planck-Institut fur extraterrestrische Physik: IX/29 ZINCHENKO I.: J/A+A/288/601 Zinchenko I.: J/A+A/361/1079 ZINCHENKO I.: J/A+AS/111/95 Zinchenko I.: J/A+AS/124/385 J/A+AS/132/211 J/A+AS/133/337 ZinK E. C.: J/ApJS/121/171

Zinn R.: J/A+A/466/181 J/AJ/110/652 J/AJ/112/1054 J/AJ/118/1671 J/AJ/127/1158 J/AJ/127/840

J/AJ/132/714 Zinnecker H.: J/A+A/312/439 J/A+A/333/619 J/A+A/347/841 J/A+A/356/541 J/A+A/359/227 J/A+A/381/941 J/A+A/392/239 J/A+A/409/147 J/A+A/422/129 J/A+A/430/137 J/A+A/440/1061 J/A+A/458/461 J/A+AS/137/305 J/AJ/121/1040 J/AJ/122/866 J/AJ/124/404

ZINNECKER H.: J/MNRAS/272/391 Zirbel E.L.: J/ApJ/448/521 Zirm A.: J/AJ/128/1990 J/AJ/131/143 J/ApJ/595/685 J/ApJS/139/369 J/ApJS/155/271 J/ApJS/155/73 Zissell R.: J/other/IBVS/5526 Zitelli V.: J/A+A/456/467 Ziznovsky J.: J/A+A/324/965 J/A+A/416/669 ZIZNOVSKY J. TARASOV A.E.: J/A+AS/107/403 Zoccali: J/A+A/448/983

Zoccali M.: J/A+A/391/945 J/A+A/429/607 J/A+A/430/603 J/A+A/432/851 J/A+A/435/657

J/A+A/465/799 ZOLA S.: J/A+A/285/459 Zola S.: J/A+A/368/932 J/A+A/374/164 Zonnchen J.: J/A+AS/133/129 Zoonematkermani S.: J/ApJS/74/181 J/ApJS/91/347 Zorec J.: J/A+A/378/861 J/A+A/385/986 J/A+A/393/897 J/A+A/409/275 J/A+A/440/305 J/A+A/451/1053

J/A+A/462/683 J/A+A/463/671 J/A+AS/129/289 Zornoza J.D.: J/other/APh/26.282 Zou W.Z.: J/MNRAS/362/1189 Zou Z.: J/AJ/117/2757 J/AJ/127/2579 VII/168 Zou Z.L.: J/A+AS/127/521 J/A+AS/132/181 Zou Z.-L.: J/ApJ/510/659 Zou Z.L.: J/ApJS/115/185 Zsoldos E.: J/A+A/275/484 Zuber J.: J/AJ/119/2801 Zucca E.: J/A+A/337/17 J/A+A/342/1 J/A+A/403/857 J/A+A/419/71 J/A+A/428/1043 J/A+A/442/841

J/A+A/457/79 J/A+A/463/519 J/A+AS/130/323 J/AN/316/39 J/MNRAS/296/599 J/MNRAS/312/540 J/MNRAS/320/387

Zuccolo R.: J/A+AS/124/425 Zucker: J/A+A/404/775 Zucker S.: J/A+A/375/L27 J/A+A/426/695 Zuckerman B.: J/A+A/275/163 J/ApJ/555/932 J/ApJS/161/394 Zukowski E.L.H.: J/A+AS/135/571 Zumach W.: II/144 Zumach w.: II/145 ZUMACH W.: J/AJ/102/395 Zurek D.: J/AJ/117/206 J/MNRAS/357/645 Zurek D.R.: J/ApJ/625/156 Zurita A.: J/A+A/354/823 J/A+A/466/905 Zurita J.: J/ApJS/170/175 Zverev M.S.: I/42 ZVEREV M.S.: I/43 Zverev M.S.: I/44 I/82 Zwaan C.: J/A+A/325/1115 Zwaan M.A.: J/AJ/124/1954 J/AJ/128/16 J/ApJS/165/307 J/MNRAS/337/641 Zwaan: VIII/73 Zwarthoed G.A.A.: J/A+AS/116/257 Zweil A.: V/107 ZWICKY F.: VII/190 VII/49 Zwicky F.: VII/4A Zwintz K.: J/A+A/462/157 J/MNRAS/357/345 Zwitter T.: J/A+A/383/188 J/A+A/417/1055 J/A+A/434/1107 J/A+A/462/795 J/ApJS/145/199 Zylberajch S.: J/A+A/303/137 J/A+A/454/185 ZYLBERAJCH S.: J/A+AS/109/447 Zylberajch S. The EROS collaboration: J/A+A/389/149